A Complete Bibliography of Lecture Notes in Computer Science (1973–date)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 5.27

Title word cross-reference

3 [KvG98, KK09a, Pol01]. << [JSdSA05]. >> [JSdSA05]. · [ETG02]. p [CKPB07].

-box [CKPB07]. -calculus [Boh75].

/CARE [GDG11]. /TC [DLNU11]. /WG [SMBQ07].

1 [AMSM^{+}09a, CC93, CP98b, EHP94, EP03a, KKGT03c, PSD^{+}07, TDBP01, TDB^{+}06, Ver00a]. 1- [Str99]. 1. [BI73]. 1/WG [BBdB12]. 10 [CMM93]. 10th [Abr07, AR99, AM03, ADP09, ABDW10, AKM10a, AKM10b, ABB^{+}10a, AFG^{+}11, AJ11, AGI^{+}06, ABM05, AF06, AWA^{+}11a, AWA^{+}11b, AHMT08, AH07, AH10, ACD^{+}10, AOM07a, AOM07b, BM96, BBMB^{+}08, BC09a, BM99a, BDK10a, BR03, BSZ12, BGK12, BdB08, BP07, BS11d, BS11e, BBB07b, BCKR10, BCST99, BMS08b, BN04, BK05b, BdBK03, BTBP^{+}08, BKV10, BDK10b, BGR03, BHT97, BG05, BH05, BLV02, BJ01c, BFC^{+}09, CSPC09, CKPR11, CVTP12, CM12, CJRR07, CRW08, CCMR03, CM04a, CMM93, Coi96, CUdlC04, CY09, CTS07, Cou03, CM10, CF00b, DLV04, DF10, DFUR11, DSZ07, DR06, DB09, Dia83, DPA05, DJRZ06, DFN10, DLŠ11a, DLŠ11b, DIP12, DPMP^{+}11a, DPMP^{+}11b, DLO03, DPNM12, DA01, DL07, EJM10, EOE^{+}07, EGMW11, EL04, EHH12, EK10, EOSW07, EFW93, ELS09, FLS06]. 10th [FRS05, FB12, FMZ05, FvdHKL06, FTZ08a, FTZ08b, FTZ08c, FTZ08d, FN12a, FLR11, Fri01, FSS06a, GHJ06a, GHJ06b, GHJ06c, GBH11, GMV07, GLMP07, GWK09, Gel09, GKS12, GGS09, GVB12, GGS11, GCSR07, GTVC06, GK02, GS10, GL92, GSV05, GC81, GF97, HDC99, HLC^{+}12, HG00, HZ08, HYPY10a, HYPY10b, HV98, HW08b, ID06, ISS^{+}06, JSWhT11, JP04a, Joh03, Jón12a, Jón12b, JMO09, JJF03, JN12, KKS11a, KKS11b, KLT^{+}12, KKS11c, KKS11d, KKS11e, KKS11f, KHG01, KŽ04, KRP06, KMMT12, KPR07, KLC07, KS08b, Lau01, LK99, LZ08, Lee04, LM06b, LPC^{+}11, Li98, LTW11, LNS11, LMA08, LLTY12, Mal09, MPS06, MSS07, MOFM10, MM07, MZ04, Mei08, MKZK06, MHK05, MR02, MMM05, MTU09, MG07b, MS10e, MWK^{+}09, NR07, NR99, NdMO03, NR98, NKLC06, NS99, OW07, OZ01]. 10th [ORB^{+}09, Ong11, OR08, PGM98, PZK07, PR11, PC99, PPJRN^{+}10, Per10a, PT98, PDK^{+}10, PCG^{+}10, PW03, PP03, PB97, PCA02, Pro11, PSS98, RS97b, RMS04, RR01, RTA95, Rei95, Rol99, RS09c, RJL^{+}08, RST^{+}10b, RST^{+}10a, Sah13, San01, SL05, SKR11, Sat97, SCHS07, Sco02, SVS08, Sei07, SG10, SLL^{+}07, SKPG12, Shv06, SCR06, SB12a, SYP^{+}05, Sma05, SPO11, SSM08, SEN08, SC09b, SPB00, SXC05, SS99, Sti90, TFK11, TBKO10, TCYZ12, TG12, VW05, VV03, VKMT11, VLY09, Wal04a, WY99, Wil01a, WM03, YY09, ZYBX08, ZOM05, dSG05, dVGR04, vB97]. 11 [CGM95, Kap92]. 116-19 [LRL98]. 118-22 [OiDP98]. 11th [ÅBE12, AM97, Ama08a, ACK00, ASR^{+}07, AS98, AHJ07a, AHJ07b, AHJ07c, AM04, AS07b, AJS00, AKR03, ADP11, ABFR07, BV05a, BVO10, BBMB^{+}08, BBD^{+}10, Bac00, BKH11, BPAB07a, BPAB07b, BH10, BSN06, BAHH07, BJM09, BI11, BKL^{+}12a, BKL^{+}12b, BN91b, BTPPS09, BP04b, BFH98, Bra02, BW01, BC09b, BHLvdP07, BMC08, BHH08, BZ08b, CRJ11a, CRJ11b, CC05, CP06a, CCPP12, CJRR07, CPC^{+}99, CSR08, CTYG10, CTS10, CC97a, CCMR05, CY11b, CK84, CG09, CGK^{+}12, CGM95, CMP07, CCSW12, Cra08, COB06, CM05b, COB^{+}08, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, DSB10, DGST09, DF11, DZ03, DPT12, DPA07, DVM^{+}11, DD07, Din04, DLGJ10, DGG^{+}10, DLJ12, EJM10, EK09, EM08, EMW94, FBSMS08, FFRS05, FR11, FS99a, FI08, FV09, FHH11, FZ86]. 11th [FLS05, FTC^{+}10, GP05, Gal07, GPAD08, Gel10, GMV09a, GHP^{+}11, Gia04, GS00a, GN12c, GS09b, GJRR08, GMS06, GG10, GM04b, GW93, GN98, GDGB12, Gud08, GP09c, GM10, HP99a, HZ05, HH04, HKL07, HOD96, HBS00, HMM12, HLW08, HKT12, Hor02, IY06, IKR00, INR07, JO99, JFBH08, JSW03, JE06, JCC^{+}12, JV06, JK05b, KR10a, KPSW10, Kap92, Kat92, KS09b, Kla12, KHPS07, KKdM^{+}07a, KC00, KFMM06, KFM07, KS04c, KKD04, KP10b, Le 02, LT00, LTW06, LLPH09, LC09, LPR12, Lin14, Lio04, LKT08, Liu11, LP00, MCA08, MBG12, Mal11, MSP^{+}09, MM01, MOBS06, MBS05a, MTOK06, MT97, McC96, MH11, MAFS08a, MAFS08b, MKZK08, MdPA98, NT98, NSS03, OVG^{+}08, OS07, PHZ11, PKM04, Pal00a, Par84, PS12a, PAD08b, PMT09, Per11a, PT92, Pet02]. 11th [PH06, PA03b, PLK^{+}12, PS11b, QMW09, QLK^{+}10, QLK^{+}11, RS99b, RR03, REF^{+}06, Rol99, RS01, RM04b, Roy05, RB00, RKS^{+}12a, RKS^{+}12b, SCKR10a, SCKR10b, SJW^{+}12, ST12a, SS06a, SYY^{+}08, SHKP08, SH07, TNP^{+}07, TTF07, Wan05b, WG11, WLRL08, XCHZ11b, XCHZ11a, ZO10, ZLY07b, ZT04, dMLG03, dSDL04, dPMA98, vB05, vdHHsHD06]. 12 [Bun94, MM97]. 12.-14 [BHH74]. 126-27 [MTT98]. 12th [AK07, AKZ01, AYV^{+}11, AR04, AL01, ABK12, ADL05, AIL12, ACJ08, BM03a, BBD^{+}10, BPSP06, BBT08, Ben06, BCD05, BCP12, BC11a, BDH^{+}99, BL07a, BTBP^{+}10a, BTBP^{+}10b, BFSN10, BCC07, BHL11, Bow94, BTT12, Bra85, Bri93, BBH08, Bun94, COBV11, CP10b, CF00a, Cer07, CW09, CL06b, CPC07, CKH10, CDCJL08, CCMR06, CP99, CA10, CCD^{+}12a, CCD^{+}12b, CH91, CSAH09, CN05, DLNU11, Deg03, DIS11, DKD05, DC00, DZGS06, DJNR09, DCFY10, DZ10b, ET01a, EKT94, ES00a, ED06b, FRS08, Fie02, FHW12, FW07, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, Foo99, FS07d, GCJB09, GG09, Gau93, GT03, Gel11b, Gel11a, GKMV10, GPR^{+}12, GH05, GH06, GZR02, God05a, GJRR08, GH12b, GGS99, GGK^{+}09a, GGK^{+}09b, GAM^{+}11, GRW86, HS05a, HM12a, HZ10]. 12th [Hed03, HP06, HAB02b, HŽ07, HTMC09, Ima99, IT08, Jac07c, Jac07d, Jac07b, Jac07a, JN10, JT09, JS11a, Jul98, JY12, KKKS11, Kap10, KLM^{+}07, KJB09, KPP08, KK11e, KS07b, KKMN07, KKH07, Kul09, KMS10, Kut98, LC06, LN01a, LH10, LTÅ^{+}11, LM08a, Leu03, LPH09, LHJ08a, LHJ08b, LHJ08c, MT09c, MS10a, MY06, MSY08, MSW07, MM09b, MS03b, MP95a, May01, MCA^{+}07, MN98, MR08, Mid01, MKZK10a, MKZK10b, MK01, MM97, MDPQA09, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, NWL06, NNPW12, OW09, Pac05, Par98, Par09, PMT10, PR05, PP12b, Per12a, PB09, PTM^{+}11, PGD07, PT06a, PRR05, RT12, RO00a, RK12, RAM05, Rol98, RMK07, SCH^{+}09a, SYMA09, Sch11, SS09b, Shy92, SQL10, SR11, SV05, Tsu08, VGM^{+}09, VW08, WC95, WLO^{+}11, WB00, WSTI08, WPD04, XSA^{+}12a]. 12th [XSA^{+}12b, YHR^{+}09a, YHR^{+}09b, YYI^{+}09, YWRS11, ZPT^{+}06, ZBH12, dA09, dRF01, dS11a, vdHS10]. 13 [Fos99]. 13th [AH00, ARS09, ABK^{+}09, AT02b, AM02, BG93a, Bat09, Bes07, BG03b, BY07, BTKP^{+}11, BDDM12, BK96, BP05, BM02b, BJKK02, BL05c, BD11, Bru04, BTMI11, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CBNR02, CCD^{+}06a, CCD^{+}06b, Che11, CCdS03, CM03b, CCMR07, CJK88, CF09b, CBF01, CFS06, CD11, Dat09, DN88, DSS11, DC12, DR11, DPV11, DLW12, DN09, DJJ12, DG09b, DJNR09, DGN01, DPT08, DB07, DFW^{+}11, Due04, EMT02a, EAES^{+}10, FKB02, Fia99, FPCK11, FvdTDG12, FG06, FR99a, FK12b, Fos99, FS08, Fre01, GN12a, GLW^{+}12, GY08, GNR07, Gel12a, Gel12b, GKMV11, GK95, GPB02, GH07b, GMV09b, GKMN10, GRP10, Gue99, HOC07, HRZK02, HCT02, Ham00, HMS09, HKP08, HN06, HV06, HSP^{+}12, HKP11, HY03, HKH10, IR08, Jac09c, Jac09d, Jac09a, Jac09b]. 13th [JL08, Jay99, Jen92, JSK10, JP09a, JNPV10a, JNPV10b, JNPV10c, JP11a, KG12, KSS08, KS03c, KBP07, Kim05b, KWCW06a, KWCW06b, KWCW06c, KS07a, KV08, Kot86, KNP06, KR12, Kur07, KŚHM11, LY12, Lin07, LXH^{+}12, LPR08, LPA00, Lou94, MT04a, MT02a, MS96b, MMG10, MMG^{+}01, MKPZ12a, MKPZ12b, MTWD06, MDPQA12a, MDPQA12b, MSS08, NSM07, NP10, OSR^{+}02, OFR09, Ong10, PDMFF12, PLC11, PRH10a, PRH10b, PHH10, PvS07, PT11, PR96, QQ11, QSWL11, RN11, RRM06, RPBP06, Rob06, RL11b, RV05, Rol99, RT10, RSK08, SPH07, Sch04b, SSJR10, Shy93, Smi06, SHKP10, Sti94, SS10, SDM05, TR12, TKCH09, Tis02, Val06, VRHJ09a, VRHJ09b, WB09, WKD08, XCR11, Yi06, YCB12, dMB05, dCHM12, vdHP12]. 14 [BS01b]. 14th [ABMM07, AMW07, ABB^{+}11, ACJ09, APBP07, AL03b, AT12, AMT03, AE06, BN09a, BYCC03, Bar03, BCE09, BPV04, Ber10, BCM00, Ber00, Bir07, BTPP^{+}12, Bod05, BL10b, BE07, BJ01b, BN09b, BS01b, BAdVR08, BL02b, BK04b, CKS04, CHD07, CFGN11, CIT11, CC12, CY12, CCMR09, CS00b, CSTD08, CD12, De 07a, DV07b, Des94, DD10, Die03, ES10, EHM11, Eta05, FMP11a, FMP11b, FMP11c, FS09b, GLSY10, GJT03, GKMV12, GO11, GR12, GSS09, GJRR11, HM11a, HWP01, HW01, Her00, Hof11, HMMW10, HW08a, IKO03, IDMY08a, IDMY08b, Jac11a, Jac11b, Jac11c, Jac11d, JMK09, KPK05, KKMW07, KW07a, Kim12, KjHGG11, KK09b, KMR12, LZL11, LJR^{+}10, LO11, LGBT03, LdQ07, LN03b, LLMR09, LGM11, LMM10, LMB08, MPK04, MB82, Man09, MRŠ03, Mar93, McC97, MNS06]. 14th [MVA01, MK96, MSS04, NF07, Nie03, OKS02, Ong05, Ott87, PM08, PR08, Pap95, PA03a, PMWO02, Pie08, PZ07b, PS12c, RR08, RDPMA^{+}11a, RDPMA^{+}11b, RBG09, RM97b, RS12, RZ12, Sag05, SS11a, SNE08, Sav06, SHØ06, SSJR10, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, SWJX12, SFN^{+}11, SM11b, Sio10, Sir12, SBM11, SM01, Stu08, SCB01, TR12, Thi94, WCK11, ZYRP10a, ZYRP10b, ZRTS03, ZYDL09, ZBY07, dGEK11, dO98]. 15 [FHP03, KK98, RKS89]. 15th [ABHT08, AV08, ATM09, ALR11, AP11, AHW07, Arn90, AJN12, ACR^{+}08, AKS09, ADGM12a, ADGM12b, ADGM12c, BS11a, BSC12, BBA12, BDCM04, Bir12, BBKM11, BZ83, BC10b, BEG04, BRP09, CMT02, CBF09, CR01a, CVV08, CS10, CCMR10, CS12a, CS02a, Cop95, CMS08b, Dan12, DS09b, DFP^{+}12, DS11c, DKOZ05b, DKOZ05a, DG97, EM03, EBT11, EP07, Ési06, FBM12, FT04a, FSV09, FHP03, FS10, FKBA12, Fri01, GTT04, GLP12, GY04, Gen09, GN12b, GH09, GJRR11, GF12a, GP08b, GPT10, GW11b, GJRS12, HMRT05, HMP08, HA02, HSP^{+}12, Hil06, HdQ08, HW03, HNQ08, HRM^{+}10, HCS11a, HCS11b, HSMPA09, HS03, KK98, KILW10a, KILW10b, KS02d, Knu01, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KKC09a, KKC09b, KP09, KR10b, KP05, LKD08, LS88, LR05a, MJ04, Mat09, ML11, MS02a]. 15th [MS00c, MSK^{+}12, MMK98, MCER10, MZ06b, Nag90, Ngo09, Nyb08, OO12, OS09, PVV05, PPH07, PT06b, RWSS12, RA12, RV10, SPBP08, SW04, SMD04, SHGB05, Ses06, SF08b, SEL09, Sma97, SHKP12, SS06c, SH10, Tha96, Thi95, TMP04, Val94, Wel01, YMY^{+}10, dGN12, dLZ12, vO04]. 16 [Emb98, Gan99]. 16th [AK10, AD05, Abe10, ABFH09, AP04b, ADW05, AMNT06, ACP05, AD11, ADCR89, BW03b, BR02, BTZ^{+}10, BMCCM11, Bra02, CJ11, CDCW09, CV96, Cho12, CMMR11, CHA03b, CV10, Coh10, CVÉ07, DGT10a, DGT10b, DD95, De 07b, DRDKE11, DD06, Dun09, DH07b, EH08, EG97, Emb98, EM10, ERMS06, FILP06, GCB^{+}05, GE08, Gan99, GF03, Gie05, GHMT05, GH04, GJRS12, HM08, JRSN09, dCJB12, KTH09, Ker12, Kob96, KHL10, KSDO06b, KSDO06a, KO07b, KSTP99, KŽ10b, Lau12, LLC09a, LLC09b, Mag02, MF11a, MF11b, Mal02, MT99a, Moh91, MHW12, MOTN07, MMM11, Ohl93, Oli07, OKdQ09, PS09a, PCH^{+}06, PH11a, Păs09, PS04b, Pue07, RCC12, Rau04a, RV11, RWD09, SM11a, SS11b, SHJ09, SK11, Sch12, SW03a, SS05b, SJT05, TCHB12b, TCHB12a, Tar91, TS10]. 16th [TP09, UWB^{+}06, VAM06, WP10, WPH03, XCd03, XYZU11, YI12, YKU11a, YKU11b, dGN12]. 17 [LJ00, McA00]. 17./18 [Sch74]. 17th [AL11a, ABFK10, ABdJS10, AMGJ12, AMRŚ08, Asa06, ACD^{+}10, AS07c, BM03a, BH06a, BLS06, BL04a, BS92, BF11, BR96a, BGS11, BRST08, BIMW12, BL07b, BKW06, BP12b, BS11f, BGH06b, CG12, Car03, CL10, CP10c, Cod01, CM11, DdQ10, Dro08, Ehr98, ELM05, EG10, ER05, FV10, FMO05, FS09a, Fic03, FYM12, FD11, FSS06b, HK92, Hen08, HK11, HI10, HMPZ09, IL01, JNR11a, JNR11b, JC12, Kal97, KM08a, KGRD10, KD05, Kin08, KCG09, LW11, LTL^{+}11a, LTL^{+}11b, Lee11, gLPZ^{+}12a, gLPZ^{+}12b, Leh11, LV06, LRL98, LJ00, McA00, MR12c, NS81, OYA04, PF05, Pat90, PSE10, Pfe06, RS97a, Rao92, RT00, RDK09, SHW09, STS04, SBG04, SvdMOP06, SP12, SMS12, TG04, VGB11, WY04, WMB10b, WMB10a, YYH^{+}12, dSADM07a, dSADM07b, vdPW10, Sad13]. 18 [Paz97, Rea01]. 18-97 [Paz97]. 18th [AS11a, ABCWM99, ABBFM11, AE05, AKR^{+}12, AR98, AM05, ABRS06, AB97, Baa07, BJ06, BL05b, BdQ11, BM10, BV12, BS93, BC11c, BAH09, CV07, Cas09a, CGS07, CDND11, CFL10, DFUR11, DKC10, DP06, Esc10, FKP99, FR01b, FS07c, FK12a, GM12, GW09, GK09, GM11b, GSS11, GMV12, Gue04, Han09, HHJZ12, HZB07, HM05b, HST07, Jou11, KPS12, KL05, KKdM^{+}07b, KY11, Kra98, KNK08a, KNK08b, LZK11a, LZK11b, LZK11c, MZ07a, MT04a, May93, Mil12, MV12a, MAA91, NDF94, OST11, RRBE09, Rea01, RD12a, SMH^{+}12, SHGB05, SCR06, ST12c, Stu02, SM09, SV12, TN03, Tok07, UDF06, Vor02, WRP07, WS12, Yah11, YŞ03, ŽL08, ZJ05, dBM10a, dBM10b, dMS09, tTVH^{+}12]. 19 [ENP02]. 1972 [SW75]. 1974 [Bli75, yF75, Hac75, KG74, Loe74, Man75, Rob74]. 1975 [Bec75, Boh75, Bra75, HS76, Nic75]. 1976 [BH77, Cow77, Maz76]. 1977 [Gru77, Kar77, SS77]. 1978 [AB78, Chi79, CER79, LB78, TPB79, Win78, Yao82]. 1979 [Bec79, Bla80, Bra80, Eng81, Kah79, Mau79, Ng79, Tob80, Wei79, YK82]. 1980 [BK80, Enc80, Jon80, MB82, MM82, Nol81, Rob80, Sal83, vLdB80]. 1981 [Deu81, EK81b, FHM82, Gec81, GC81, Koz82, Sta82]. 1982 [Bet83, Cal82, ENR83, Lov82, NS82, SS83]. 1983 [BZ83, CK82, Dia83, Kar83, KC84, vH83]. 1984 [BCI85, BRW85, CK84, Par84, PR84, RRG85, vdV84]. 1985 [BJ86a, Bra85, Ehr85a, Ehr85b, GJ86, Jou85a, Jou85b, Nic86, Pic86, Roz86, Wad86, Wil86]. 1986 [BRR87a, BRR87b, CDW86, FK87, GRW86, KBD87, Sel86, Wad87]. 1987 [DMWW88, FTF88, Nos88, Roz87, ST88]. 1988 [CJK88, For88, FTF89, Jos88, Roz88, dBdRR89, vZvdV89]. 1989 [FTF91, Roz90, Stu90]. 1990 [Roz91, Sti90]. 1991 [PGV93, RRP91]. 1992 [Roz92]. 1993 [Roz93, SBC^{+}94]. 1996 [Ant97, CRW97, RV98]. 1996-2000 [DLFP02]. 1997 [AKL^{+}99, MZ98, PWZ98]. 1998 [HK98a, HMT99, HSTU99]. 1999 [Apt00, Gir99, Mal99, MSY00, Pol99a, Wal99a]. 19th [ABFV12, AD06, APPE09a, APPE09b, ACW07, AF02, APHA11, ADK04, Baa03, BKW08, Bla05, CBGCCZ12, Can12, CHZ08, CS05b, CCMS11, CM09c, DH07a, DV10, De 10a, De 10b, DKK08, DH11, DS98, DP12, ENP02, EH12, FL08, Fra05b, FGTZ08, Gor10, GSW06, Gup10, HNF08, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, JW05, JCC^{+}12, KCMO11, KvdM00, KOS07, KRSR11, Kuc11, Kui92, LLS00, LM06a, LS06c, Mar11c, MS12c, MvL10, Ong05, OdQ12, Pal03, PVTG07, PRR94, RW11, RRR99, RBW95, ShK06, Tis94, TBD12, Vor08, Wie99, ZdC08, vBC08, vKS12, vL94b]. 1st [BMK08, Bis88, BBR88, DD07, DJS91, FL09, HPP88, KJ88, KL88, NS87, Sam76, Sch92, SM91, Sim92, ZSM^{+}09, dS02].

2

[CC95, CP98c, EP03b, HSS96, Pol86, Ver00b]. 2/WG [MW08]. 20 [JLY03]. 20.-22.3.1974 [Lei74]. 200 [BO00a]. 2000 [AH00, AKU01, ACK00, AM00, AJS00, AJ01a, Bac00, BdBC01, BDPS02, Bel00, Ber00, Bha00, BBH01, BP01, BLKW01, BO00b, BGP00, BB00a, BNS00, BKM01, BDGK00, BCD00, BW00, BH00, BMS01, BSS00, BE01, BJ01d, Cag00, CSCO00, CLMP02, CRL01, Cas01, CL01, CD00a, CKO01, CW00, Chr00b, CW01, CS00b, CR01b, Con00, CRS00, Cup00, CF00b, DM01a, DM01b, DMW00, Dec00, Del00, DDJ00, DC00, Du00, DA01, Dwa00, Dyc00, ES00a, ET01b, ES00b, FL00, FV01, FR00, FIAP01, Fra01, GM00, GS01, GFV00, GS00a, GPV00, GS00b, GSS00, GKOT00, GW00, Hal00, Ham00, HWP01, Har01, HG00, HBS00, HS00, Her00, Heu00, HJW00, Hor00, HJ00, IKR00, IZ00, JK00, Jos00, KMT00, KP00, KM04, KR00a, KR00b, KK00, KP01b]. 2000 [KvdM00, KM00, Kow01, Kro00, LLS00, LSW01, Lau01, Lau00, LT00, LBP00, LPH00, Llo00, LPA00, LP00, LZ00, LK00, Mai00, MMM01, Mar01, MF01, Mil00, MS00b, MK01, MS00c, NL01, NW01, Nar01b, NS00, NR00, OAdGBP00, Oka00, Oli00, OTZ01, Omi00, PP01, PDH01, Pal00b, PV00, Pat00, POS00, Pol00, PC00, PR00, Pre00, RS00, RO00a, Ras01, RT00, RGW01, RSL01, RH01, Rol00, RO00b, Rus00, ST00, Sch01a, SvS00, Smo00, SKP00, SZ00, SMMG01, SPB00, ST01, SV01, Tah00, TSG00, TLC00, TG00, Tiu00, VJKT00, VKV00, Ver00a, Ver00b, VWY01, WB00, Wat00, WMR00, Whi00, Won01, Yak00, Yas00, YP01, Zan00, ZS00, ZY01, ZKZ00, vLWH^{+}00]. ’2001 [BBZ01, AKZ01, Abr01, Akm01, AFM01, Amb01, AL01, Ari02, AM02, BBE01, BST01, Bar03, Bat01, BGV01, BMP01, BB01a, BND01, BS01a, BK02, Boe01, BU01, BKS01, Bos01, BJ01b, Boy01, BL01, BJ01c, BFMS01, BS01c, CGS01, CR01a, CWL01, Cod01, Col02, CK01, CBF01, CS01a, Cou01, CS01b, CMM01, CH01, DF01a, DH01, DF01b, DG01, DAB01, DMR01, DST01, DS01a, DSV01, Die03, DGN01, DBH01, ET01a, EV01, EFT01, Esp01, FR01a, FR01b, FMW01, FZJ01, FG04a, Fre01, Fri01, Fro02, GM01, Gel01, GG02a, GJRT01, GK01, Gom01, GLN01, GSP01, GNTF02, HMV01, HW01, HK01, Hof01, HM01, Hus01, IL01, JS01, Jen01, Joh01, JS02, KSHS01, KWA01, KS01, Kil01, Kim01b, Kim01a, Kim02, KR01, KZ01, Knu01, KP01a, KNP01]. 2001 [KMSS02, KRMM^{+}02, KYT01, KU01, KRS02, KJS01, LRS03, LSW02, LN01a, Lee01, LMW01, LN01b, LYhL^{+}01, Liu01, Lor01a, Lor01b, LMŠT01, Mal01, MEK03, MM01, MY01, MR01, MŠKL02, MB01, Mat01, Mat02, May01, MT02b, Mey01, Mid01, Mil01, MP01a, MP01b, MPV01, MVA01, Mon01, MDBF01, Mos01, Mue01, MJL02, Nac01, NV01a, NV01b, OZ01, OPT01, OSvL01, PR01, Per01, Pet02, Pfi01, Pic01, PG01, QOZ01, QBA01, Ram01, RD01, RRR01, RS01, Sak01, San02, San01, SS01, SPK01, She01, SLC01, Sil01, SMK01, Ska01, SLL01, Smi01, Ste01, SM01, SCB01, Syv02, Tah01, TLF01, TWYhL01, TNN^{+}01, Tuc01, VM01, VY01, Vog01, WS02, Wal01, WYL01, WW03, Wel01, WvdH03, Wes01, Wil01a, WL03, WHS01, WWC02, WDPW02, YM01, Zho01]. 2001 [Zit01, dRF01, dRS01, dS02, vSN01, vdBV01, vdL02]. 2001/2002 [FG04a]. 2002 [AO02, AFvZ02, AT02a, AS02a, AK03, AMU03, AH02, APSZ02, AF02, AFS02, ABM02, ABJ02, AMCL02, AT02b, AT02c, Arm02, BV02, BW03a, BCT02a, BCT02b, BS02a, BTQ02, BPR02, Ber02b, Bha02, Bis02, BR02, Bla03b, BS02b, BJN03, BMO02, BM02a, BM02b, Bra02, BLV02, Bre03, BJKK02, BL02b, BB02, BCF^{+}02, BLPW02, BD03, BHM^{+}02, BLS02, BPS02, CAD^{+}02, Cag02, CBC^{+}02, CDP02, CZ02, CGP02, CBNR02, CM03a, CG02, Cha02a, CCL02, CUDL02, Cha03a, CLY02, CCH02, CM03b, Cho02a, Cho02b, CGP03, Coe02, CPvdH03, CS02a, CGS02, CEKR02, Cor02, CP02, DR02a, DO02, DB02, DG02, DQBZ02, DR02b, Dim03, DS03, DK02a, DK02b, DDS02, Dor02, DKR02, EF02, ES02, EMT02a, EMT02b, EL02a, ETG02, EL02b, FHJ^{+}02, FBKS03, FV03, FRS02, FKB02, Fie02]. 2002 [Fle02, FG04a, FCLL02, FLUV02, Fos02, GRT02, GM03, GW03, GOW03, GZR02, GKW02, GPB02, GK02, GAB02, GCC^{+}02b, GCC^{+}02a, GP02, GAR03, GG02b, HP02a, HRZK02, HG02, HCT02, HTW02, HA02, HM03, HS02, HP02b, HM02, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, HRT^{+}03, HAB02b, HH02, Hor02, HRA02, IZ02, IS02, IT03, JLV02, JKL02, JM02, JJP^{+}02, JJ03, JHC02, KKP02, KWA02, KR02a, KMD02, KS02c, KS02d, KOS02, Knu02, KH02, KC02, KMTU03, KR02c, Kuc̆02, LO02, LSS02, Le 02, LV02, LL03, LhK03, Leu03, LSE02, LFK^{+}02, LMP04, LHK02, Mag02, Mal02, MT02a, MS03b, MS02a, MTP^{+}02, MS02b, MSW02, MKZ02, MR02, MF02, MK03b, MS02c, NP02, NE02, Nür03, NH03, OSR^{+}02, OKS02, Oka03, OYY03, PDHS03, Par02, Pat02, Pau03, PV02]. 2002 [PA03a, PS02b, PH02, PBGK03, PK03, Pet03, PTZ03, PMWO02, PKSS02, Pre02, PCA02, RM02, Ric02, RK02, RV02, SKS02, SVS02, SMB03, SUW02, Sco02, ST02, She03, SBD03, STDH02c, STDH02a, STDH02b, SKP02, SMK02, STTP02, Stu02, Sur02, Tis02, TBJ02, TG02, TL02, THR03, Van02, VS02, WK02, WVD02, WTM^{+}02, Win04, WPH03, WKB04, Yak02, YAF^{+}02, Yun02, ZSD03, Zhe02, ZJS^{+}02, dBBGdR03, dBBC02]. 2002-2005 [dSOSR06]. 2003 [ACvLdR03, AKK04, AS03, AH03, AFKL04, AL03a, AFL03, AMT03, AGM03, AJRS03, ABD03, BM03a, BLPW03, BYCC03, BST03, BCD^{+}03, BS03a, BP03, BI03, BBKW03, BLB^{+}03, BB03, BG03b, Bih03, BdBK03, Bla03a, BL03, Bod03, BHU03, Bon03, BGR03, BS03b, BM03c, BZ03, BK04b, BCR03, Bru04, BHM03, BKB03, BBC04, BFOY04, BKO03, CCS04, CDV03, CLM03, CV04a, CT03, Car03, CY03, Che03a, Che03b, CH04, Chi03, CHA03b, CUdlC04, CGGO03, Cou03, CDM04, Cro03, Cue03, DW03a, DD03, DDE04, Deg03, DSS03, Des02, DZ03, Dig04, Din03, DKB03, DTW03, DW03b, EM03, EP03a, EP03b, EW03, EF03, FPRAW04, FD03, FRS03, FSM03, Fic03, Fon03, FLL03, FMY04, GH03, GMS03, GJT03, GF03, GMV03, GT03, Gel03, GMO03, GHSW04, GT04, Gor03, GPS03, GTY03]. 2003 [GAR04, GKN03, GY03, Had03, HV03, Hed03, Hic04, Hof03b, HW03, HMG03, HY03, HH03, HS03, ID03, JSO04, JSW03, Jen03, JP03, Joh03, JM03, JJF03, Joy03, KS03a, Kah03, KCR04, KS03b, Kay03, KS03c, KN03, KOOL03, KS04b, KHW03, KLPP03, Kra03, KWT03, KKGT03c, KKGT03a, KKGT03b, KCO04, Lai03, LN04a, LS04a, LGTB03, LGBT03, LB03, LOST04, LCF^{+}04, LL04, LN03b, Lio04, LM03, LCY03, LK04, Mag03, MP03a, Mal03, MS03a, MRŠ03, Mar03, MA03a, MVMP^{+}04, MM03, MP03b, MTS^{+}03, MT^{+}03, MSS03a, MDP04, NNS03, NdMO03, NZ03, Nie03, NT03, ND04, NSS03, Oho03, OPP04, OS03a, Oqu03, OWPY03, PHJ03b, PHJ03a, PR03a, PH04, PCPS03, PBH03, PR03b, PW03, PU04, PPS03, Pez03, PNB04, PS03, PBK03, PP03, PA03b, Pri03a, Pri03b, QGZ03]. 2003 [Rai03, RR03, RABR03, RS03, Ros03, RSK03, SNS03, SRS03, Sar03, SKMT04, SW03b, Seb03, SZC^{+}03, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, SG03, SLLS03, SDH03, SCKR03, TN03, TBH03, Tra04a, THT03, VV03, VJAA03, Vei03, VC03, VKJ03, Vos03, WSS04, WKP03, Wan03, WZ03, WvdH03, Wha03, Wii04, Wim03, WR03, XCd03, YŞ03, ZG03, ZYH03, ZJXC03, ZNCF03, ZZO03, Zuc03, dBBG04, dLWC03, dMLG03, vEDA04, vdAB03, vdAHW03, vdL04]. 2003-2005 [KLT^{+}07]. 2004 [AAW04, AR04, AL04, AP04b, AP04a, ADMH04, AB04, ABT04, ADK04, BDZ04, BHH04b, BHH04c, BMU^{+}09, BR04, BCL03, BPV04, BMR^{+}04, BEP04, BDCM04, BCP^{+}04, BI04, BK04a, BEGP04, BEG04, BGKS04, BD04, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BDFS04, CC04, CMZL04, CvSW04, Chi04, CHAL04, CM04a, CR04b, CM04b, CSSW04, DSZ04, DSB04, DFM04, DB04, DPB^{+}04a, DPW04, DKLS04, DH04, DLdS04, DFT04, DFD04, Due04, EB04, EFM04, Ekl04, EW04, FC04, FMC04, FM04, FKK04, FCD^{+}04, FCLG04, GTT04, GSH^{+}04, HK04, IMW04, JYZ04, JKRR04, Jay04, JP04a, JPXS04a, JPXS04b, JGXC04, JK04, JVA04, JQ04, Jue04, KG04, KMW04, KS04a, KWW04, KV04, KGL04, KP04, KOL^{+}04, KŽ04, KOK04, KFW04, Koz04, KS04c, KR04b, LGK^{+}04c, LLMT04, LLWL04, LRG04, Leo04, LVP04, LWF04]. 2004 [LN03a, LN04b, Luo04, MPK04, MJ04, ML04, MT04a, MEAC04, MT04b, MPvH04, MTC04, MH04, MKZB04, MKR^{+}04, MAFSS04, MSS04, MSUB04, Nao04, NHJ04c, NCBT04, NdG04, NBK04, Nor04, Ode04, OS04, Oka04, OWM04, OYA04, PKM04, PD04, PS04b, PP04, Rah04, RUL04, Rau04b, RM04a, RR04b, RKW04, RSTZ04, SW04, SMD04, SRG04, Sch04a, SLH04, STT04, SCH04c, SKP04, SKK04, SL03, TG04, TN04, Tra04b, TSKGB04, VPBŠ03, VMBMN04, Wal04b, Wan04, WPV04, WMS04, WPD04, Yak04, YGGJ04, YEY04, YJ04, YX04, YWG04a, YWG04b, ZJ04a, ZZ04a, ZGY04, ZJ04b, ZSP^{+}04, ZT04, dFEN04, dSDL04, vO04]. 2004-2006 [BMU^{+}09]. 2004/2005 [AGM05]. 2005 [DG05a, GWS05, HK05b, Roy05, Ytr06]. 2005-2007 [TNGK08, dCdF^{+}08]. 2005-2008 [CKOS09]. 2005/2006 [Col07, MBJ^{+}07]. 2006 [BMU^{+}09, IMK06, ISS^{+}06]. 2006-2008 [DF09]. 2006-2009 [LT11a]. 2006/2007 [AG07, PS07a]. 2007 [EP07, MVC07, Mül10]. 2007/2008 [Mül10]. 2007/2008/2009 [ABG09b]. 2008 [GSYZ10, Mül10]. 2008-2010 [MN12a]. 2009 [BCG^{+}11, CLLD09, GCJB09, ZSP10]. 2010 [LRH10]. 2010.Proceedings [GHK^{+}10, RMB10]. 2011 [ADP11, BBKM11]. 2012 [HSP^{+}12]. 2012.Proceedings [YYS^{+}12]. 20th [AGP08, AW12, Alp11, AH03, Bar11, BG09, BL08, Bel00, BKW09a, BZ09, BM11b, BG07b, CFLR12, CMSA12, dRCVT11, CF10, DL06a, DL04a, DDI10a, DDI10b, DDI10c, Dol06, DDI09, EF12, Ési06, FKK04, FKM09, FL11, GLZZ09, GH96, GM08b, JLY03, KP00, KL07b, KD99, Kno11, KW07b, Kuc11, KU09, KJS01, LKC93, LM08b, LTT07, MDY07, MMDN12, MST95, MK12, Nie05, OA07, OT07, SB07, SC11, SHUH08, Tho06, Tre09, Vog01, WH95, YGGÖ05, YAKM^{+}12, vLS11]. 21 [AJ01b, WM04]. 21st [AS94, AP10, ABM02, AGCP^{+}11, BA07, B^{+}12a, Ber08, BM09a, BHN97, BHQ10a, BUHS08, CCP10a, CCP10b, DH04, DH07b, Ern07, GG05a, HMV01, HDGK11a, HDGK11b, HSVZ10, IS11a, JY14, KM11, Kil01, Kir96, MSWZ06, MMT08, MS11a, MTNL12, Nag95, NBGA08, NS00, NBM09, O’B12, Pel07, PS96a, Pfe07, PPM09, RLKHL12, Sei12, SHW09, SMK02, Vid12, WM04, Win09, WZ08, vEGW09, RLKHL12]. 21th [Ama08b, B^{+}12a, BHQ10b, GL10a, LSY^{+}06]. 223-25 [LK00]. 22nd [AS02a, AFL03, AiNOW11, Atl08, BSW95b, BMSHW09, BNUW09, CHCA09, CK01, DD05, ET06, FG95, GJ09, GPCL10, GL10b, GM11a, HM11b, HLSZ11a, HLSZ11b, HG98, IS11b, JNS05a, KM08a, KK11a, KSUZ11, LMA09, NDFMS97, NL09, PV02, Per10b, PSM10, PR97, Sch09b, SLLS03, SH11, Tau08, TCJ10, VDÉ^{+}12b, VDÉ^{+}12a, Vit08, Yun02]. 23rd [ACL^{+}04, BH06c, BSS11, Bon03, B^{+}12b, BCC99, CMCS11, CFB11, DN07, Dro09, DT06b, EL02b, FK10, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GH12a, GQ11, GK09, GV09, HLW04, JKB96, KS12a, Kei09, KHW03, KU12, Li11a, LSTZ12a, LSTZ12b, MCSN10, MM96, Moh97, MR11, MSKY10, RF01, WZ11, ZBG98]. 24 [RR89]. 242 [RMV96]. 24th [AB12, BFBS11, BL11, CK07, Cra05, CGvR02, D’H10b, DV10, DGMS97, DKM^{+}05, FJ10, Fra04, HS98c, KPW99, LM04, LS10b, MS12a, MMO^{+}11a, MMO^{+}11b, NPS06, PJ97, RFBW12, TW07, WR11, WGD05, dFEN04, dlBP08, vG02, vdAB03]. 25 [GV08]. 25th [BV92, CR04b, EOR06, Gór06, GJS08, HRB12, HW09, JKL02, JDAW12, KI12, LSW98, Li11b, Mez11, MK03a, NR00, PPCS12, Pel11, RS06, Rov98, Seb03, Sho05, SBH^{+}04, Vau06, Wan05a, WNE99]. 263 [CRS00, KS03b, Smi01]. 264 [SCKR03]. 26th [Agu12, AKG06, BW00, CD05, CBCGA12, Dwo06, GKN03, MT04b, NPPDG06, Nao07, Nob12, PSST07, PTB99, RBGS04, SO07, Sex09, SPK01, vWN99]. 27th [AP07, BFP04, BL01, DV07a, DR02b, DT06a, HS08a, HJW00, KP15, KSH05, LSYO08, LFL05, Mac12, Men07, Sma08, WMR00]. 28- [Kow01]. 28th [BRS09, FGB11, FMNS06, Fur05, Jou09, KY07b, Kuc̆02, LCD^{+}09, LMR^{+}06, OSvL01, PR01, RV03, SHYEF08, Wag08]. 29th [ACR07, Bod03, FKS07, Gil10, GP02, Hal09, HSJ07, LLPH09, PSS^{+}10a, Sch10b, WTM^{+}02, vHV08]. 2nd [ABB12, ACN^{+}07, AGS91, BG91a, BB90, BMS04a, Ber02b, BMJ88, Bod91, BBR89, Bra75, CG88, DSS91, DT89, Dit88, DPP97, DW07b, ENR83, Fil91, Gan88, GML^{+}08a, GHR89, GM89, GK90, GS91, GPV88, Ham89, HN12, HL91, KO91, LB78, Mau89, Meh85, MO05b, Nos88, PRdRW09, Pol86, RA00, Rei89, SCR06, TBE08, vL88, vdV84, vzGIK08].

3 [Cal86, CCP96, HMP99, RTB02, vdL01].

3. [Bra73, SF74]. 30th [BLPW03, BD11, CSA^{+}12a, CSA^{+}12b, FBV11, FW09, HZ10, HBE07, HNW04, JS05, JDL11, MOP^{+}08, Pat11, Rab10, Rig08, VPBŠ03]. 31- [CJ11, HNP12]. 31st [ACR12, DBB^{+}08a, DNS09, DKLS04, KU06, Kra05, LP10a, OD12a, PJ12, Rog11, VBCBS05]. 31th [BBMSD09]. 32nd [CIM^{+}05, Fom06, GR12, GRK^{+}10, GHK^{+}10, KP11, KK07, MHA09, SNC12b, WTP^{+}06]. 33rd [BKM07, BPSW06a, BPSW06b, CFG^{+}11, DBHS10, HP12, MF11c, OT08, vLIvdH^{+}07]. 34th [ACJT07, BE11, BYdVZ^{+}12, BEFP08, GKB^{+}08, KN09b, OF15, PPBL12]. 35th [ADG^{+}08a, ADG^{+}08b, GK12, HK10b, NKM^{+}09, PH10]. 36th [AMSM^{+}09a, AMSM^{+}09b, MS11b, PPBL12, Thi10, vLMP^{+}10]. 375th [Van89]. 37th [AGK^{+}10, ČGH^{+}11, KK11d, RSW12a]. 38th [AHS11a, AHS11b, BFG^{+}12, GSLM12]. 39th [CMPW12a, CMPW12b]. 3D [MT09b]. 3DPH [MT09b]. 3rd [BR89, BC93, Bli75, BBR90, BMYZ91, BP09b, Cal86, CZG^{+}08, CW89, Der89, DA08, DL81, Ehr87, ELNM08, GANACL11, GMUW12, Jon90, LS92, LS89, LMŠT01, MWD06, Mai88, MW91, MMP03, MVN86, NS90, PDG92, Rei88, SILS05, SCKR03, TDG91, VZ99, vF91, vdABCC05].

4 [BTF02, Est95]. 4. [Sie75]. 48th [Vit10]. 49th [BV11]. 4th [AO02, AP09, ALS^{+}11, AN97, AW01, Akl95, AN95, ABMS06, AF11, AJB09, AYBH^{+}06, ADM^{+}08, APS06, Apo93, AMPT09, AS12, AJ94, Arm02, ABD08, AL12, AB06, Avg12, BG10a, BKH11, BC95, BE06, BS11c, BPCS06, BMPM09, BCKK05, BBM^{+}08, BLS12, BBAB^{+}07, BTQ03, BHP95, BBP^{+}08a, BBP^{+}08b, BPR08, BDL08, Bec75, BE08, BS03a, Ber96, BR97, Bes93, BND01, Bha05, BW10b, Bih97, Bil97, BH92, BBZ01, Bla03a, BBH01, BK02, BB10, BGU94, BC05, BRU06a, BDS04, BJ01a, BB06, BTF02, BGP00, Boo91, BB00a, BDDS07, Bor91, BH02, Bos00a, BAP^{+}05, BVNW87, BCF06, BHJY07, BW92, BDW97, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BD03, BKO04, CV11, yCCZ07, CLKL09, CDV03, CDP^{+}05, CGL^{+}09, CV08, COMTLB12, Cas01, CW94a, CR05a, CD07, CA08, CY03]. 4th [CKK^{+}09, Che03b, CMS08a, CF09a, CM09a, CH09, CQSS09, CW00, CFG96, CKM^{+}09, CLB11, CBT07, CWO^{+}09, CAdB01, CKC11, CPP07, CDS09, DGSM11a, DL06b, DD97, DB02, DF01b, DMR01, DFMV10, DQBZ02, DL95a, DH10a, DWML05, DJS06, DEM96, DSV01, DMWW88, DFN97, DL04b, DH06, DRS05, DSL08, DTW03, DBB^{+}04, DH08, DFS06, Dyc94, EV96, EL07, EH95, EKR91, EO94, EHRT08, EdS08, ELN^{+}10, EK07, ES92, FL00, FV05, FSdlPT11a, FSdlPT11b, FFL^{+}12, FS99b, FTAZ06, FDL10, FSN06, FJL97, FX08, Fon00, FG96, FP05, FKM07, Fra01, FG10, FdS12, FT94, FLMK06, GP09a, GANACBR10, GANACBdCdV12, GAL12, GJ93, GGK06a, GKC^{+}12, Gel03, GdATM05, GRB^{+}07, GM03, GKS^{+}07, GMBM05, GMO03, Gir99, GL05, GJRT01, Goe98, GK01, GHSY06, GPV00, GP12, GN05, GZ06]. 4th [GL03, GF12b, GAN05, GM05, GRR05, GPP09, GYD^{+}06, Hac75, HENR03, HT11, HS07, HV03, HS94, HK10a, HM03, HBN06, HC11, HAB98, Hof01, HK96, Hol08, HS98b, HM01, HNP12, HKNW07, HKV10, HO98, HK98b, HWLJ08b, HWLJ08a, Hus01, ICSY07, IMY^{+}07, IFL^{+}12, IS12c, JPBT05, JOD98, JM07, JP07a, Jan93, JS01, JBJ94, JNHJ10a, JNHJ10b, JLS11a, JLS11b, Jeu98, JJ03, Jha10, JM03, JK04, JLW07b, JP07b, JP96, JMP12, JMO10, Kåg98, KSG^{+}09, KN09a, KKP02, KWA02, KC07, KR02a, KS03b, KMD02, KP92, Kim01a, Kim02, KS06, KKKN05, Kit88, KN03, KK00, KP01a, Kob06, KFW04, Koh06a, KRH09, KB92, KR02c, KMYJ08, KP10c, Kur10, KMKP11, KFfKL11, KBC05, LPT07, LV07a, LSW02, LLCL05, Lau00, LMW01, LCY10, LWDM12, LG99]. 4th [LLC^{+}05, LA94, LLM^{+}08, LMWY04, Liu01, LCY03, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LBDW09, Lom93, LSF07, LD05b, LD05a, Lou92, Luo07, MP98, MRvM07, Mal97b, MMBND07, MS03a, MS10b, MS10c, MVW09, Mar05a, MSZ09, MVCM10, MS05, MM89, Mat01, MBP^{+}11, MS99b, Mil01, MS06b, Mos01, MS02c, MNW98, NP02, NW01, Ng93, NHJ12a, NHJ12b, NV01a, Nik05, NCJK08, NP11b, O’H98, ODP^{+}07, OPP04, OK96, ÖRH12, OZ11, PR07, PDH01, PCN12a, PCN12b, PCN12c, PZ08, Pat02, PF06, PBH03, Per05, PI05, Per11c, PSD12, PGBK04, PVP05, PSN95, PSNS99, PT99, PKSS02, PBRB08, PT91a, PT91b, PR98b, RF10, RPR08, RFZ03, Rau11, RMFG06, RB10, SBV^{+}11, SS07a, SFS11, SYNC11, STP^{+}10, Sch95, ST11, SH08, STL11, SS94, Sch06, SVV12, SE10, SP08, SBC^{+}94, SRW98]. 4th [SA12, SC98, SP93, SH09, SMCM07, SA97, SDH03, Ste98, SW97, Ste07d, Ste07b, Ste07c, STTP02, SD10, SWMM06, SC05, SSZ12, TM09, TMS12, TLZ^{+}08, TLC00, Tom92, TNY07, TKHS05, TSKGB04, Vad07, VY93, Van97a, VW07, VA07, Vet11, VMBMN04, VMB^{+}12, Vor93, Wad86, WAS06, Wei79, WLP^{+}09, Whi00, Wim03, WTGA05, WFB12, WR03, Win04, Win11, WCBD08, WD10a, WD10b, WDPW02, XLKZ12, XYM^{+}07, YN06, Yan11a, YLWY06, YHV12, YY01, YLL09, ZBW94, ZSSM03, ZEL04, ZL98, ZYB06, ZF05, ZKZ00, ZJS^{+}02, ZVU99, ZHP12, Zuc03, dBBGdR06, vLS91, vdL02, ASP01].

5 [Est95, HP89]. 50 [LIBP07]. 50th

[FN12b, LIBP07]. 5th

[AM10, Abr01, ADDL08, ANS04a, ANS04b, ANS04c, AD98, AH02, ADR07, ANF08, AS06, ASLM09, ABD^{+}99, AMCL02, ABK12, ABG96, AT02c, AZ11, AJ94, Asa99, AB03, ADS09, BO00a, BCBF08, BSvRW08, BCT02a, Ban93, BLW09, BYLJ11, BPP91, BSW95a, BP04a, BBB^{+}07a, BMSMS04, BGP11, BBP04, BBP^{+}09b, BBP^{+}09a, BHHK08, BWAM09, BCC^{+}10, BR92, BC10a, BG11a, BMS08b, BB05b, BC06, BLB^{+}03, BG11b, Bha07, BF00, BSBM05, BP01, BBC^{+}08, BGU95, BPS06, BMM07, BRW93, BG10b, BI04, BC07, Bor92, BF06, BF10, BMYZ95, BDN08, BKM01, Boy95, BCD00, BC11b, BFvdM10, BHLvdP07, BFMS01, BZ03, BV05b, BW09, BP93, BRR97, BBB06, BW97, BT06, BLVV07, BN98a, BN98b, BFKO05, CHKL10, CDP^{+}06, CK08, CV04a, CV04b, CE01, CM95, CCJ07, CG02, CCL02, Cha05b]. 5th [CJ08b, CWL01, CFK^{+}12, Chi03, CCBW11, CCLR98, Cle99, Coe11, Col02, CUdlC04, CLS07a, CS01a, CR73a, CR73b, CRS10, CW88, Cou93a, CMB^{+}08, CR05b, CGB11, CG94, CDA^{+}06, CK93, CMQ96, CEER96, DW03a, DKP05, Dam97, Dam99b, DM06, DVVA08, DD03, DSRW08, DH01, DCFR12, DY06a, DGG^{+}11a, DIMV11, DGM09, Deh97, DP11, Deu81, DSM03, DKLT05, DCM82, Di 97, DTMH10, Dim03, DK02a, DK02b, Dol09, DW03b, DGB^{+}06, DZ94, DLM07, Dwa00, DHSH96, DS07, EKS10, EB04, EFG^{+}09, ERRS10, EBSS98, EML^{+}12, ETG02, Feh11, Fer98, FMS10, FK02, FHY08, FMN99, Fos02, Fri94, FLT07, FKL08, Gab13, GP11, GANACL11, GFV00, Gel04, GRG06, GLP99, GMP06, GH10, GHSW04, GOZ05, GP98, GJRT01, GW11a, GZ09, GPPW08, GCM08, GV95, GLM97, GR06b, Gra05]. 5th [GHH^{+}06a, GW06b, HG02, HPTV05, HHMT12, HR96, HRC06, HKI08, HS76, HW01, Ho12a, Ho12b, HM09, HW06b, HMG03, HM04, HdL12, HYP^{+}09, HJL^{+}09a, HJL^{+}09b, Hug91, HP89, HL00, HMCL06, Jaf99, JLV02, JS97, JCC^{+}12, JHC02, JP08, JP94, JDMB08, JRR03, KS08a, KM11, KP06, KMW03, KJR05, Kar94, KR04a, KL96, Kat99, KK11b, KY07a, KWC^{+}12, KSW05, KM04, Kir93, KZ01, KH02, KS10a, KSHM09, KU01, KRS02, KR06, KA05, KL02, KS07c, LLT07, LSS02, LSW03, LM98b, LL03, LSL^{+}09, LHŚ11, LT05, Len97, Leo09, Lev98, LLT^{+}04, LWF04, LV05, LEP^{+}10, LSS^{+}04, LFK^{+}02, LY05, LMW06, LG95, LLL06, LWYN09, LK04, LHK02, LG05, LS06b, LKK10, Luo08, MMD05, Mah04, Mai90, Mal99, MST12, MNZ09, MMR07, MSS12, MS12b, MVL11, MTV09, MPPJ^{+}05, Maz76, MMMO96]. 5th [MGH09, MV12b, MMK08, Mon96, ML95, MSR99, MdSC09, MC05, NP02, NYHO07, NZZ11, NMPM08, NP12b, NKPH08, NE02, NSFB10, ONC^{+}11, ONPR07, OGM05, OYM07, Oli00, OPS11, PRV11, PHR09, PT09, PDHS03, PGA^{+}05, PKM^{+}10, PA09, PFD^{+}10, PW09, PF08, Per07b, PT08b, PCG^{+}05b, Pet03, PPS03, PMKG07, Pfe94, Pic01, PK00, PHOK10, PMC06, PBS08, PG09, Pro96, Pup99, QGZ03, Qui98, RG05, Raj02, RS10, RRU05, RVY05, RS00, RT07, RR98a, RTA95, Reu97, Ric02, RKW04, RBC93, RCS10, RSvdP05, RJS^{+}08, RG11, SGPP06, SL12, SZR^{+}08, ST88, San94, SB95, SEG09, SV97, Ser12, SDH05, Sha07, SJ06, SKP02, SAUL11, SMF^{+}11, SMMG01, SHL08, SL03, Ste09b, Ste09d, Ste09c, SCKR03, SRT06, STY^{+}07, SRL12a, SRL12b, SRL12c, SZT^{+}08a, SZT^{+}08b, SNSR06]. 5th [SvASD05a, SvASD05b, SvASD05c, Syv02, TM99, TFE^{+}09, TKSN11, TJ08, TG02, TN08, TSK92, TGAH11, Tri98, TNGK08, THT03, Vau98, VJAA03, VU05, Ver08, VSH11, Wad87, WSS04, WKP03, WZD11, WZ09, WK02, WL04, WVD02, Wim04, WSGA06, WZA10, WN96, WHR10, WKS^{+}10, WDPW04, XW11, X^{+}12b, XL11, YEY04, YGE12, YT99, YJKK06, YKMJ09, YP01, YGL^{+}10, Zam08, ZZP^{+}10, ZHLW12, ZJXC03, ZZO03, ZSP^{+}04, dBBGdR07, dRS01, vDKS11, vdHCD07, vdL04, BK80].

6 [BD98, CH98, Fra00, Mor89a]. 6.3/7.3 [HHG11]. 6/TC [DSB10, DC12]. 6/WG [KD05, SHUH08, UDF06]. 60th [AG09, ACG^{+}07, AAN09, ACÇ^{+}12, BDFM12, CLKK07, CS12b, DKN12, DKS12, EMO10, HS05b, KK11c, KMO^{+}05, LLM09, Mah04, MRU12, MvOvRdV05, Pal09, PQR11, RSB07, SPG10, SW10]. 64th [Der04]. 65th [BS11b, BKT12, DDM08, ELS^{+}10, FJM06, GKKW09, Hei12, HNR05, Len04, Nac12]. 68 [Bra76]. 6th [ACN^{+}07, AFvZ02, AK97, AT02a, ACD^{+}07, ASE^{+}07, AvAS06, AvASD06a, AvASD06b, AvASD06c, Alo05, ABR09, ALPH10, ABJ02, Ang97, AI98, AJRS03, ATC^{+}09, APP12, ABW10, AB81, ADL09, BSvRW09, BS09, Ban94, BKS99, BLO^{+}07, BHKR11, BMH05, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BHRU09, BGKY08, BB01a, BJM09, BLKP08, Bha95, BG12b, Bil98, BD98, BBRM98, BH77, Bla03b, BBD^{+}09, BPT06, BM02a, BKS05, Bor93, BGPR10, BR07a, BKM08a, BDGN07, BM03c, Bra09a, Bra07, BGHS05, BD04, BGR08, BJ12b, BS12, BM06, Bue04, BR07c, BDLMM12, BS06d, BFKO06, CFI06, CNX05, CBG^{+}10, CFZ10, CZF10, CP10a, CT06, CJ08a, CGP02, CCK^{+}08, CGBZ08, CHCM11, CLY02, CC09a, Che10, CKLN10, CEQ^{+}11, CHAL04, Chr99, CB06, CH98, Coe98, CR01b, CKOB11, CKW11a, CKW11b, CF99, Cou05, CS01b, CP02]. 6th [CK82, CLBP11, Cup00, Dal92, Dam99b, DG06a, Dar97, DSDS04, DK90, DSB04, DMS04, DG07a, DDHP05, DFM04, Dec00, DMV12, DG10, Deh99, Dem07, Des02, DLN10, DOS99, DGR06, Dil94, DLM99, DBB^{+}08b, Dov05, Du00, DRS08, EEKR00, EEKR12, EFT01, EM06, EP03a, EP03b, EMT02b, EV07, EKM06, FRvKS10, FHJ^{+}02, FKP^{+}05, FWY05, FC04, Fer08, FLMG10, FD93, FLP12a, FHLP09, FB09, Fra00, FT04b, FLL03, FPRS06, Fri04, FKLT08, FTF88, GU95, Gal97a, GHM05, GVM99, GANACBdCdV12, GVT08, Gel05, GM07a, GCK06, GT04, GJN10, Gor03, GR06a, GS00b, GMS12, GYA^{+}06, GTY03, Gru98, Gru77, GPL99, Gyi96, Hal94, HS12b, HHM97, HHLS08, HG96, HS00, HKP10, HU08, HMW^{+}08, HP94, HA00, HPS09, HS95c, HK05c, HGZ05, Hof03b, Hsi95, HRA02, HH03, HZBF10, HZGZ10]. 6th [HLFS12, IPRO08, IT03, IN11, JSV11, JM10, JSZ95, Jay04, JS97, JL00, JKN^{+}12, JM11b, Joh97, Jon00, JP09b, Jos00, JDMB10, JS94, JQ04, KMW04, KC09, KKMW07, KS01, Ker97, KSB10, KGG^{+}11, KOS02, Knu99a, KRRBV07, KW07c, KPK^{+}10, KPS09, KS12c, KHI06, KH11c, KO07a, KKL12, KS12d, KSGT10, KH12, KV11, KDWW10, LJS08, LYL11, LM00, LS10a, LS95, LB03, LHKŚ12, LHH10, LM09, LLLL08, LOS05, LCF^{+}04, LL04, Lin12, LP10b, LMCH12, LMW08, LYhL^{+}01, LL06, LWRdR11, LQO04, LV07b, Luo09, MRN07, MMV06, MP03a, Mal01, MBT^{+}03, MN02a, MM08, MD04, MA03a, MP10b, MR12a, MJR04, MPT10, MM03, MRPM10, MO08, MH04, MA11, MMU96, MCSY07, MS00b, MNLM06, MC89, Mor89a, MMC05, MNS95, MSX10, MAR05b, Mue01, Mug97, MC98, MZ06a, NNS03, NKN^{+}05]. 6th [Nie96, NRWI09, NR08, Oka00, OMMSP06, ODV08, OPKR05, PGM98, PMÓ06, PMH09, PR84, PF10, PpWR96, Per06, Per09b, PM96, PGG^{+}06, PU04, Pez03, PMD97, PRJ95, PP11, PBW05, RiN12, RMDS10, Ram08, RZ91, RA00, RST09, RK06, Rei09, Reu99, RT95, RYG11, RS09b, RV02, REPH06, RSK03, Sab10, SNČ^{+}12a, SHM^{+}10, Sch10a, SSRA98, SF11, SDR^{+}00, SZ92, SZC^{+}03, SKK08, SMEGA10, SCH04c, Smi12, SKY06, Sta95, Ste11e, Ste11f, Ste11d, Ste11c, SWB03, SLCT11, Sur02, TWQ^{+}12, TDWG11, TT06b, TNI09, TOL08, TAS05, TNGK08, VM01, VWH06, Ver00a, Ver00b, VKJ03, VV07, WLC^{+}06, WW03, Wet92, Whi98, WBW94, WSG07, WJP07, WDM06, XBM12, Yal10, YRWS11, YAKMJ10, YLLZ04, YHZ09a, YHZ09b, YHZ09c, YZCW07, ZPTI09, ZS05b, Zui99, dBBGdR08, dCVK07, dSTP11, dS02, tCZ07]. 6th [vHH09, vSN01, vdHXMW08, vdMS11, von96, BLKW01, Lov82].

7 [Gac02, RTB02]. 70th [ADM11, BDR10, CRS11, HL05, JLW07a, JPR04, KMPR04]. ’74 [BHH74]. 75th [JL11, MM09a]. ’79 [Ng79]. 7th [APFV09, APR^{+}06, AMS06, AF97, ACK^{+}07, ABBC10, AR10, AP04a, AJRS03, AAH^{+}10b, AAH^{+}10a, Asa96, Aus93, ACR^{+}08, BSMM04, BBdB12, BBvRL10, BJ10, BR12a, BCD^{+}03, BDZ04, BW11, BL12, BFH07, BCS09, BHH04b, BHH04c, BK05a, BHJKPG06, BL06a, BS02a, BPW06, BBP^{+}11a, BBP^{+}11b, BD12, BB99, BIRT05, BCRB^{+}11, BDG^{+}12b, BLJ08, BMS04a, BH07, BSTL07, BD97b, BTPPS05, BBR^{+}10, BS02b, BCC^{+}12, BGM94, BG07a, BGL09, Boy01, BvdHPP09, BBT10, BB11, Bro92, BC04b, BS06c, CD10, CCF^{+}08, CL11, CK10a, CK10b, CBS^{+}12, CS03, CZ02, CL01, CDGW10, Cea76a, Cea76b, CM03a, CGKM11, CCM^{+}04, CWZ11b, CEK11, CL90, Chr00a, CMS11, CM00, Con00, CYBF06, CSA^{+}12a, CSA^{+}12b, CBW99, CHK06, CvH07, CSRS08, CPR12, CBBS09, CSSW04, CT99, DFL^{+}10, DO02, DG04, DPLW08, DJAP07, DEKR09]. 7th [DPC^{+}07, DG09a, DEG11, DST01, Del00, DGSN11, DBM09, DGWB09, DH94, DDK11, DFPS06, DKP00, DMW06, DBD^{+}10b, ELS12, Ehr91, ECB06, EN06, ENO12, EF03, Esp01, ES00b, FR01a, FKK09, FR09, FP08, FKBP11, FHW08, FSP09, Fuc98, FG04b, FTF89, Gac02, Gan96, GD10, Gel06, GM08a, GGHN97, GKT09, GS09a, GH07a, GMUW12, GKW02, Hae99, HKR07, Hal00, HY12, Han98, HK94, Har07, HPV00, HJ12, HT05, HS95a, HC05, HAB98, HPP06, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, HKLP12, HM96, HS11, HPRS06, HM05a, Hu09, HGBF12, HGGG12, HGPH12, HDLM09, ILR95, IR07, IST07, IS02, JP05a, JM11a, JKRR04, JM02, Jen01, JRPP07, JP10, JS02, JVA04, JP12b, JK06, KMTW03, KS04a, KLZ07b, KLS10, KOL^{+}04, KMSB11, KKH09, KOOL03, KS04b, KKM^{+}11, KBHSW06]. 7th [KC02, Kos98, KBČ^{+}12, KHS07, Koz04, Kra03, Kra99, KLFK10, KJ10, KS10b, LK10, LPOA^{+}06, LB05, LD97, LGTB03, LhK03, LYY07, LN09, LYBU10, LVP04, LW09, LZS^{+}08, LKZ^{+}08, LCB12, LNW^{+}12, LN03a, LS06a, LM03, LMW10, LLO08, LR09, LMT10, LRH10, LN04b, LNSS11, LCG97, LS12b, Luo10, MCKB06, MPV12, MSB12, Mal03, MPF95, ML04, MY01, MMP07, MRP11, McG08, MFG12, MNPW07, MC94, MAB^{+}02, Mic10, MR07b, MR97b, MKB^{+}09, MS00c, MS96c, MPS04, MV06, NNS06a, NNS06b, NZS^{+}06, NO11, Nes99, NZ03, Nie93, NL99, OROD07, OV09, PZ07a, PKKG06, PHJ03b, PHJ03a, PP01, PVW10, PS04a, Pal00b, PH04, PV00, PC05, PT08a, PRK12, PMA^{+}07, PFM12, Per07a, Per11b, PCG^{+}07, PVTG07, PN93, PMDK00, Pin95, PRG09, PVV10, PdA^{+}11, Por98, PM09]. 7th [QMLW05, RS05, RK93, RLKHL12, RRR01, Reu01, RL11a, RSTZ04, Sak01, SSO^{+}11, Sch93b, Sch01a, SW03a, SvS00, SA96, SSD^{+}08, SvADS07a, SvADS07b, SvADS07c, SvADS07d, SKSA12, SKB10, Slo99, SKP04, ST07, SDP97, SZ05, SS09d, ST01, SZ11, SA04, SKR^{+}10, TLC06, TKCW11a, TKCW11b, TW12, TT05, TBF^{+}09, TND10, TEB09, TNGK08, Urz05, VKV00, VP96, VBH07, VH04, V^{+}06, WT96, Wal99a, Wal01, Wal04b, Wan01, WZJX12, WT07, WR07, WMS04, Wha03, WSF08, Win78, Wol95, WDKW08, XKGS12, XBS^{+}10, XZCL07, Yam08, YAKMJ11, YKL06, YLC^{+}10, Zan00, ZZ04a, ZTY^{+}05, ZLK10a, ZLK10b, ZSO99, ZCGH11, ZYRH06, dBBM09, dSK12, vSH06, vdBHOS09, vdHIP11, MA03b, MA03c, Sho84, Wri03].

8 [Sak91, San02, WD00]. 8.-19

[BRR87a, BRR87b]. 8.4/8.9 [TQYX11]. 8.8/11.2 [DFPS06, GLIC10, GS08, Pro11]. 8.9/TC [QBY^{+}12]. ’83 [vH83]. 84 [BC85]. ’85 [BJ86a, Buc85, Cav85, Ehr85b, Wad86, Wil86]. 85th [ADR08]. ’86 [Wad87, Aus86, Odl87, TS87]. 8652 [TDBP01]. 8652/1995 [TDB^{+}06, TDBP01]. ’87 [Bjo87, FTF88, BHCL87, BBR88, CP88, Dav89, EEK^{+}87, EKLM87, GS88, NS87, Nos88, Pom88]. ’88 [BMJ88, FTF89, WLS89, BB90, BBR89, DN88, Gan88, Gia89, GN88, Gol90, Gün88a, GPV88, KSW90, SCM88, vL89, MLM90]. ’89 [Bra90, DO89a, FTF91, MT89, PMD90, BB89, BF89, BBR90, Buc90, DSS89, DO89a, DO89b, GM89, Jan89, Mau89, Nag90, QV90, RCA90, Der89]. 8th [AG01, AB11, AJRB09, ARSB09, ABG09a, APS08, AS11b, AH96, Amb01, AYB^{+}09, ATO^{+}09, AGS^{+}11b, AGS^{+}11a, AH97a, AÖ12, AOSV08, ABRZ06, AP83, BM03a, BW05a, BMK08, BGLT05, BDE12, BGV01, BBP^{+}12a, BBP^{+}12b, Bec79, BB95, BDIR07a, BDIR07b, BPR02, BCG^{+}07, BDG04, BKR09, BDG^{+}12b, BDW08, BF09, BDZ08, BKH^{+}09, BRL08, BCP99, BLSS11, BC93, BVAT12, BTPPS06, BMSW12, BL03, BR12b, Bv97, BLvdH10, BG08, BK04a, BC10c, BEGP04, BZ08a, BKLS11, BDV95, BL06c, BW05b, BAWH00, BvADS08c, BvADS08a, BvADS08b, Bud91, BFK^{+}07, CPS05, CYM^{+}11, CIW06, CCD04, CCD^{+}09, CCJS07, CFB^{+}05, CT03, CP12, CDCW09, Cer05, CP11, CC07, CMMR12, CH10, CDPT06, CJRT05, CCAS12, Chr01, CW06, CVV12, CDN12, Com97, CDSS07, CMV96, CGGO03, CP07, CDL12, CC97b]. 8th [CD01, Cou01, CL08b, CDT11, DSZ04, DH05, DG06b, DBC^{+}10, DSS03, DKN^{+}11, DEM96, DPP09, DW10, DLW^{+}07, DBB^{+}12b, DGJ09, DF06, DG05b, Dwy01, EK12, EKP^{+}07, EPVW10, ÉEMS04, FRS02, Fle99, Fle02, FKS10, FR99a, FNPW05, FSTS09, FHP11, FLP11, FTF91, GM00, GMV05, GMO09, GMS03, GRT02, Gel07, GKT10, GG02a, GLM^{+}08, Giu98, GM06a, God05b, GW06a, GM06b, GS08, HP02a, Had03, HO03, HW06a, HT97, Har09, HK98a, HOD96, HL04, HCS^{+}05, HDBM11, Hen06, Hig05, HSKB09, HB11, Hon01, HSH08, HNS^{+}12, HYMZ11, HSB^{+}96, HJBF12, HMJG12, HMT10c, IZ02, ID03, Ibr95, IAC06, IAL^{+}07, IK12, Ish11, Jäh99, JKRR04, JSO11, JJP^{+}02, JFT11, Joh01, JP11b, Jue04, KEDW07, KC11a, KC11b, Kap08, KS02c, KTC^{+}05, KYL07, KM04, KS93, Klu97, KOK04, KHSM^{+}08, KW10, KS96]. 8th [KS10c, LH11, LBNF08, LW95, LCHS08, LBM^{+}10, LIJ97, Leu08, LM97, LMW12, LN01b, LH06, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, LP08, LS09, Luo11, MWD06, MWYF10, Mal05, MEK03, Mar01, MW11, MMP05, Mat02, MW97, MPR12, MW09, MT08c, MM10, MT02b, MKZ02, MPPU11, MPV01, MF02, MTC^{+}08, MT05b, MDBF01, MSH05, MA07c, NSS^{+}07, ND09, NHJ04a, NHJ04b, NHJ04c, NV01b, NQL06, Ode04, OLL10, OT11, Ölv10, OSV11, OS03a, OKBMB10, PT95, Pag98, PM04a, PM04b, PM04c, PM04d, PAD^{+}08a, Pat00, PS02b, Per08, Per12b, PJM^{+}08, PNPP12, PRG10, PLI08, PW07, QBA01, RR12, RZ94, RG03, Roc07, RS03, RBGG06, RGLA09, Rus00, RTZŻ06, RSW12b, SST00, SS08a, SNS03, SKS^{+}06, Sak91, SRS03, Sar03, Sas05, Sch04a, SNRG11, SWAF95, SLB05, She01, SG03, SL99b]. 8th [SLKL09, SRY07, SS04, SSv94, SHUH08, Swi99, TdLdOQ08, TV94, Tho11, TSS11, TT06a, TP94, TNYL11, Ued10, UPB07, UMN^{+}08, Uus06, Vah09, VZ06, VTK^{+}12, Van02, VAR12, VY01, Vau05, VG12, VD12, WSJT09, WDKK07, WK06, WVDS05, WD00, WDKW10a, WDKW10b, YKKZ07a, YKKZ07b, YBL^{+}04, YYS^{+}12, YTC^{+}07, YJKSW10, YAKM^{+}12, ZGY04, Zhe02, ZLDB05, ZLS^{+}06, ZY10, dBBHL10, vHC08, vdBV01, vdPS08, Sie86].

9 [Est99, KDS01, MMR91]. ’90

[BHL90, CC91, Ham91, vE91, AK90, Arn90, BK90, BTT90, Bor91, BMYZ91, Dam91, Fil91, Jon90, MV91, Mio90, NS90, Sch92, SP90, SSB90, MR91, AFK90]. 90-VAPP [Bur90]. 90’s [SdDK91, Bla90, KN91]. ’91 [AM91b, AM91a, DSS91, Fei91, GS91, vF91, AvLR91a, AvLR91b, AP93, Ame91a, ABS91, BG91a, BPP91, BG91b, BS92, BN91a, Bor92, Bud91, CH91, Dav91, DFK91, DKM91, DER94, Her91, HL91, IRM93, KJ91, LS92, MW91, PMD92, PT91a, PT91b, Pri91, RZ91, RB91, TWN92, TSK92, vF91, Boo91, Kod91]. ’92 [Chr91, Dos93, NT92, PW92a, Rue93, BC92, BH92, Bor93, Bri93, BW92, CW94a, Cle92, CFR92, Der92, DGR92, ES92, FNY93, Fit93, FGM92, Iba92, Jan92, KB92, LM93, Lou92, Mad92, May93, Nak92, NU92, PT92, PDG92, Rao92, San92, SZ93, SZ92, Sim92, Tom92, VY93, Vor92a, vP93, BR92, Ohl93, Pet92]. 92-VAPP [BCRT92]. ’93 [BRW93, Cou93b, Deh93, DN93, Gau93, GT93, LG94, AZ96, BN94, BG93a, BUG93, Bes93, BG93b, BGM94, BP93, CM95, CTT93, CKM93, CK93, DN93, Dyc94, EKT94, Esi93, FD93, FO95, Fri93, GJ93, GNR94, Hee94, Hel94, Jan93, JS94, KS93, Lam96, Len93, Lom93, MP93, Mio93, MCP93, Ng93, Nie93, OA93, PMD94, Ral94, RK93, RBC93, Sch93b, SBC^{+}94, SP93, Sti94, Tor93, Vor93, WAR94b, WL93, vL94b, Bra93, Kir93]. ’94 [AJ94, DPP95, FT94, GO94, KK95, TT95, AYB95, BU94, BGU94, BS94, BCP94, Bor94, BM94, BMYZ95, CNY95, De 95, Des94, Dil94, DW94, DZ94, DNS95, Ekl94a, Ekl94b, FT94, Fri94, HM94, Hal94, HS94, HP94, IY95, JBJ94, JP94, Jou94, Kar94, KMN94, Le 94, Lei95, LA94, Lou94, MPP94, MST95, NDB94, NM94, Nie94, PT95, Pfe94, PSN95, RZ94, San94, SS94, SSv94, Ste94, TC95, TV94, TP94, War94a, WBW94, ZBW94, vL94a, BdR94, CO94, DL95a, HKM95, LR94, NDF94, WJ95]. 94-VAPP [BV94]. ’95 [Aya95, BB95, Chi95, GQ95, MR95, Myc95, WS96, WGL96, Ada96, Akl95, AN95, BYGP95, BSW95a, BSW95b, BHP95, BB95, BC96a, Bha95, BGU95, BH95, BDV95, Bra96, Bri95, BGY96, Bun96, CCK95, Cop95, DCP95, DMW96, DvV96, DHMM96, EH95, Ell95, FR95a, FK95b, GS95, HAM95, HP95, HR95, HS95a, HS95c, ILR95, KL95, LS95, Li96, LMV95, LG95, MPF95, MNT95, Mol95, MR95, ML95, MNS95, Nag95, Olt95, PMDA96, PRJ95, RM97a, Rei95, RT95, SB95, Sch95, Sel95, Spi95, Sta95, Vit95, WC95, WH95, Wol95, WMT96, AH97b, Bar97, Hsi95, LW95, Mal95, RBW95, TD97, VA95]. ’96 [APO96, AH96, CS96, DF96, DP97, Fer96, Fre96, HG96, Hof97, LM96, MML97, PZH96, Str96, vGH96, ABG96, Asa96, BM96, Bod96, BK96, BMP96, BC96b, BC96c, yCW96, CL96, CH96, Coi96, CMV96, DS96, DPP97, DHSH96, EEM96, FR96, FG96, FGL96, Fre96, FM96, GO96, Gae96, GW96, GPM98, HR96, HK96, JY96, KL96, KM96, Kir96, Kob96, KS96, LM96, MS96a, Mau96, MM96, MMMO96, MMU96, MS96c, Mon96, NDFMS97, Nie96, Nor97, PZ96, PS96a, PpWR96, PS96b, Pla96, RM96, Ret97, SOS96, SA96, SLPM96, SC96, Tha96, VP96, VT97, VWY97, WT96, WN96, YKF97, Gan96, GH96, MD96, Mug97, SF96]. ’97 [BCB97, BM98, Con97, DFN97, Fre98, Gae96, GF97, HHM97, HLR^{+}98, HB97, TGM97, ter97, AI97, AN97, AK97, BD97b, Bih97, Bil97, BBD97, BM98, BHT97, BR98, BRR97, BW97, CC97a, CC97b, DMO98, DH97, Deh97, Del97a, Del97b, FM97, FR97, FJL97, Fum97, Gal97b, GL97, GB97, GGHN97, GHK97, GLM97, HOQ97, HHM97, Hir97, JS97, JL97, Kal97, KK97, Ker97, KZ97, LLP99, LIJ97, LM97, Li98, Luk97, Mal97a, MMT97, MT97, MW97, Mos97, MBC^{+}97, NT98, Par98, PT97, PMD97, PB97, PR97, RS97b, RPZ97, SV97, SK97, SA97, SW97, WY98, dRLP98, BHN97, Com97, Kit98, LD97, LCB97, Mal97b, Paz97, TC98, YD98, vW97]. ’97-HOA [HHM97]. ’98 [GN98, Heu98, IZ98, Ish99b, Kåg98, Kam99, LMV98, PDH99, RH98, SKB98, Wil99, AG98, ANR99, ACL98, AI98, BM99b, Bil98, CPC^{+}99, DS98, DTF98, Ehr98, FR98, GWY99, GW98, GGS99, HDC99, Han98, HAB98, Hir98, HO98, HK98b, HS98d, JOD98, JR98, Jeu98, Jul98, KW98, KvG98, Kos98, Kra98, Kut98, LRL98, McK99, Moo01, MB98, MNW98, NS98, NK99, Nyb98, PS98a, PS98b, QSBC98, RR98a, Ric98, Rol98, Rov98, SSRA98, SCG98, Ste98, SHO99, TM99, Tri98, Vau98, WC98, ZBG98, ZQ98, dO98, AK99, COT98, FC98, HG98, HS98b, HSTU99, MdPA98, MMK98, Nip98, Pag98, Qui98, SMPU98, SC98, dPMA98, NR98, WKK98].

’99 [Bau99, BDH^{+}99, BBZ00, Fon00, GdlP99, Nes99, NL99, PSNS99, PVC^{+}99, Pol99a, Str99, Tho99, Asa99, BCP99, Bra99, CH99, CW99, DOS99, FS99a, FS99b, FR99a, FR99b, Fra99, FM99, GLP99, GM99, Gup99, HP99b, HJRS99, HL00, HS99, JO99, KP99, Leo00, LG99, LIH99, Mur99, Rol99, Ste99, TC99, Wie99, BCC99, DF99, HKB99, Ima99, Jon00, Mal99, MTO99, MT99a, NR99, PC99, Pup99, VPK00, ZZ99]. 9th [AGB^{+}11, AI06, ACL12, AdBB12, ABKT11, AL04, ANS^{+}09a, ANS^{+}09b, ARC^{+}07, ABMH08, APW06, APV09, AGL06, APM08, ACR^{+}08, ACD^{+}10, BV02, BMS^{+}08a, BMK09, BdBC01, BM05a, BMUV10, BA99, BBS07, BH06b, BPS11, Bat01, BW10a, BJP12, BP09a, BPZ09, BDSW09, BCP^{+}04, BG06, BTPPS07, BJ12a, BJ07, BNS00, BL02a, Bos00b, BH95, Bro94, BCR03, BOH^{+}12, BH11, Bun96, BHLP07, BFK^{+}08, CHM^{+}10, CLM03, CK12b, CK12c, CFM09, CD00a, CJRT05, CR04a, CH04, CvSW04, CRD08, CDY^{+}11, CCMR02, CSY09, CI98, CHD98, CCM08, CU08, CAB10, CTE^{+}06, CML^{+}10, CS02b, CLP05, CFP^{+}09, CKTW12, CC09b, Cra12, CFRR10, Cur09, DR02a, DR05, DLOS05, Del97a, Del97b, DS01a, Dem80, DBB12a, DGLT12, DJRZ06, DKU05, DKB03, DOSY05, DLW^{+}07, Dur12, DF99, EHKP05, EKF12, EO94, EKR10]. 9th [Esi93, Est99, FT11, FPP11, FC98, FD03, FRS03, Fes10, FJ92, FHKQ12, FST09, FLUV02, FCDB07, FKLY08, GGD09, GMV06, GH03, GB99, GH08, Gel08, GKST11, GHK97, GA09, GLIC10, GHC^{+}06, Gru97, HK04, HMT10b, Han07, Har11a, HMTH10, HR95, HJMP06, HWG10, HP02b, HHW01, HT06a, HCL05, HCDL09, HC09, HXC^{+}08, IS08, JS08, JP12a, JTB^{+}05, KMMN11, KLB^{+}06, KHJ05d, KT06, KSB11, KDS01, KR02b, KO11, KTB09, KLPP03, KBH03, KKDV02, KASS07, Kru08, LO02, Lai03, LNS06a, LNS06b, LS04a, LLWL04, LRG04, LBP06a, LBP06b, LBP06c, LBP06d, LS04b, LPZ08, LST^{+}12, LO10, LT11b, LMŠT01, LGS09, LSW^{+}10, Luo12, LHH11, LIH99, Mah04, Mal07, MR97a, MŠKL02, MBVTY11, MT07a, MMR91, MEB05, Mel07, Mey01, MM04, MKZB04, MMK^{+}05, MDP04, Mor06]. 9th [MV07, MJL02, NdG04, NWHR06, NRR12, NW10, Nip98, NBC^{+}06, NBM09, NSL09, NH03, eOP05, OP97, Olt95, Oqu03, PV07, PDML11, PBK^{+}10, PZYS12, PSPH^{+}10a, PSPH^{+}10b, Pat03, PY05a, Per09a, PDF^{+}09, PH12, PRG11, PW08, QIW07, QLXL07, QSBC98, RCJ^{+}08, RM96, RCT05, RL06, Ric98, RMTW11, Rol98, Ros03, RD12b, RBB^{+}06, RTZZ08, RKNV09, SKS02, SSVW12, SW09a, SRG04, SPCG07, STC04, Sd98, SH12a, Sch06, Seh97, SO09, SOS96, SCL^{+}06, SS06b, SAG10a, SAG10b, SCP^{+}12, SL99a, Ska01, Smo00, SKP06, SOP12, SEN07, SP07, STWS07, SHRM06, TYZ11, TTM10, TLJ06, Vää99, VJ11, VZM^{+}10, Wan03, WPV04, WYP12a, WYP12b, WZ03, Wat00, WR09, WCJS10, WKB04, WHS01, WANL08, WDKW12a, WDKW12b, YW06, YMHH10, Yan11b, YX04, YDKM06, ZK06, ZiTM10a, ZiTM10b, ZiTM10c]. 9th [ZSM^{+}09, vW97, vdMvdT08, vGH96, LO88].

A4MMC [VMvN12]. AAAI [HMBD09]. AAC [JPXS04b]. AAC-GEVO [JPXS04b]. AACC [MAD^{+}04]. Aachen [BS12, CS96, DG01, FSTS09, HHJ^{+}86, HR96, JKL02, JLS11a, JLS11b, KK11a, KS96, KMN94, MK01, Nag95, POP09, RST09, SLPM96, SLKL09, TW07, Wei79, vO04]. AADEBUG [Fri93]. AAECC [BL07b, BAH09, FILP06, BS01b, Cal86, CMM93, CGM95, Fos99, FHP03, HP89, MMR91, Mor89a, MM97, Pol86, Sak91]. AAECC-10 [CMM93]. AAECC-11 [CGM95]. AAECC-12 [MM97]. AAECC-13 [Fos99]. AAECC-14 [BS01b]. AAECC-15 [FHP03]. AAECC-16 [FILP06]. AAECC-17 [BL07b].

AAECC-18 [BAH09]. AAECC-2 [Pol86]. AAECC-3 [Cal86]. AAECC-5 [HP89]. AAECC-6 [Mor89a]. AAECC-8 [Sak91]. AAECC-9 [MMR91]. AAIM [ALZ11, Che10, CP06b, FX08, GZ09, KL07a, MXZ05, SLSW12]. AAIP [SKP10]. Aalborg [EP07, FT11, HSL^{+}99, JR98, LS92, LSW98, LN01a, MSP^{+}09, NZ03, SHØ06]. AAMAS [BPD06, CvRVSN12, DFPV11, DHtM^{+}12, FRA05a, FS07a, HKM^{+}07, HMBD09, JSS08, KHK^{+}08, KVMH09, LSJ06, NVSB^{+}07, PSP^{+}02, PAV^{+}10, SPON08, TT10, VKG12, FBKS03, FPRAW04]. aan [HM11b]. Aarhus [AH97a, Bri95, BFMS01, CDCJL08, Cra05, Dam91, DF01a, GM01, Jon80, Mey01, MNS95, NS82, NT98, NS00, Olt95, PDF^{+}09, RTB02, SS94, Sta82, ST12c]. Aarts [CA06]. Abdominal [YC11, YSL12, YHV12]. Aberdeen [CP02, GL92, MHK05, SMF^{+}11]. Aberystwyth [She03]. Abstract [BH10, BC93, CP07, Cou05, Ehr91, EO94, EN06, FBHW00, HOD96, JS11a, JMO09, KR12, LPZ08, RTA95, ST88, SL03, Cor02, Kle89, Mul93, She85, US90, Zuc03, Bjo80, BGR03, BBBB08, DFG^{+}12, FGK^{+}10, Geu98, GKOT00, ZT04]. Abstraction [Sch03b, CW00, KH02, MR07b, ZS05b, MO97]. ABZ [BBBB08, DFG^{+}12, FGK^{+}10]. AC [DvEF06, vEHD05]. Academgorodok [BBP93]. Academia [BP12a, SMMG01]. Academic [BF10]. Academy [MTH06a, MTH06b, MTH07a, MTH07b, MHD09, SXC04]. ACAI [LMŠT01, Nos88]. ACAI/EASSS [MŠKL02]. ACAL [KRH09, RAW07]. Acapulco [CSCO00, KCO04, MA05c]. Accelerators [SHM^{+}10]. accents [Gor02]. Access [AFP^{+}10, BMC08, CFH^{+}12, CDM04, FGK^{+}11, FKLT08, Kam99, PCG^{+}05b, PDF^{+}09, PDK^{+}10, PCG^{+}10, RG03, SBV^{+}11, SZC^{+}03, SS04, SP07, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, VBS^{+}10, CS03, Leo00, PGBK04, Ste05]. accessing [GT05, PGG^{+}06]. Account [MR12b]. Accra [DG09b]. Accurate [MT86]. ACCV [KKS11c, KKS11d, KKS11e, KKS11f, KH11a, KH11b, NNS06a, NNS06b, YKKZ07a, YKKZ07b, ZiTM10a, ZiTM10b, ZiTM10c, Li96]. ACCV’98 [CP98b, CP98c]. ACE [NRR12]. Acess [Ste07d]. ACG [vdHHsHD06, vdHS10, vdHP12]. Achievements [DGG^{+}11b, DP10b, GV08, OLL10, SHRM06, BSW12]. ACII [DGSM11a, DGSM11b, PPP07, TTP05]. ACIIDS [NLŚ10a, NLŚ10b, NKJ11a, NKJ11b, PCN12a, PCN12b, PCN12c]. ACISP [BSN06, BG05, BN09b, MSS08, PH11a, PSNS99, PGD07, SH10, SMS12, BS02a, BCD00, PS96b, SNS03, VM01, WPV04]. ACISP’97 [VPM97]. ACISP’98 [DB98a]. Acitrezza [GE08]. ACIVS [BTPPS05, BTPPS06, BTPPS07, BTBP^{+}08, BTPPS09, BTBP^{+}10a, BTBP^{+}10b, BTKP^{+}11, BTPP^{+}12]. ACL [Dig04]. ACM [Alo05, BC09a, BFMS08, BFJM09, BB00b, CC07, Fei03, GM10, IS08, Jac04, KK11e, Lib07, NL99, TDK^{+}10, vSH06, BCT02b, Bis02, DM01a, ES03, Gue01, Gur04, Hug91, JS97, MB98, San02, SC00]. ACM-L [TDK^{+}10]. ACM/IFIP/USENIX [Alo05, BC09a, CC07, GM10, IS08, Jac04, KK11e, vSH06, ES03]. ACML [ZW09]. ACN [hKA10]. ACNS [APFV09, BSZ12, BGKY08, GKS05, IKY05, KY07a, KS10a, KS12c, LT11b, ZYB06, ZY10, GSP01, GPS03, JYZ04, ZYH03]. ACoS’98 [MSRP98]. ACPC [Bos96, Vol93, Zim92, ZVU99]. Acquisition [HhKRT06, KR10a, Nes95, Paz03, RK09, RK12, Aus93, FS99a, PB97, SOS96, SSv94, Wet92]. Acre [EK81b]. ACRI [BMUV10, ECB06, SB12a, UMN^{+}08, BCT02a, SCH04c]. Across [DS08, KMSS02, RBTD04]. AcrossGrids [Dik04]. ACSAC [CPC07, JE06, SXC05, OS03a, YX04].

ACSC [KL95]. ACT [ZL96]. Action [ACÇ^{+}12, BKLM07, EFZKM07, EBAH08, EV09, EHMM09, HS12a, MWD06, Smi03, SS06e, TSE^{+}09, dS02, dSOSR04, dSOSR06, CFR92, FKKB^{+}05, Lau93b, RMV96, SK97, SZ00, Thi00a, Had94]. Actions [Fre98, vEV12]. Active [BP04a, CU08, ECV^{+}10, KP10b, MTU09, SPG10, SR11, TR12, UPB07, Dov05, ALPH10, AH98, CW07, Cip95, Cov99, HDLM09, LYhL^{+}01, LWYN09, MNW01, MCSY07, SS99, STTP02, TYNM05, WSS04, Yas00, ZCGH11]. Activities [SG07, JSdSA05]. Actors [ADM11, TP97]. Ad [ASE^{+}07, BY06, BGW06, CSJD06, DPLW08, ENO12, JWH06, Sch10a, SMCM07, WW07, ZOCJ07, CHPW05, MTW05, SC05, CSRS08, FLR11, KO07a, KR06, LPR12, NBK04, NW10, PBK03, RGLA09]. Ad-Hoc [BGW06, ZOCJ07, BY06, CSJD06, JWH06, SMCM07, CHPW05, MTW05, SC05, KO07a, KR06, NBK04, NW10, PBK03, RGLA09, CSRS08, FLR11, LPR12]. Ada [AK07, BP12b, BLW87, Chr91, Gau93, Kat92, KV08, KK09b, LS04b, Luc87, PH06, RV10, RV11, TDBP01, Tou94, VW05, Goo83, LS04b, Uhl82, Asp98, AH97b, Bar97, Bar08b, BO80, BS02b, Chr91, CS01b, Gau93, GdlP99, HM87, HB97, Kat92, PK00, RS03, Str96, TD97, TDB^{+}06, Tou94, Tou96, Uni81, UA83, Wal84]. Ada-Europe [AK07, BP12b, Chr91, Gau93, Kat92, KV08, KK09b, LS04b, PH06, RV10, RV11, VW05, LS04b, Asp98, BS02b, CS01b, Gau93, GdlP99, HB97, PK00, RS03, Str96, Tou96]. Adanced [XYZU11]. Adaptable [Paz99, PvS07]. Adaptation [DKC10, HMPZ09, MMDN12, TNGK08, AKK03, DEM81, Gor02, KKA05]. Adapting [DGSN11, PPPM10]. Adaption [KCMO11, TtHVS06, VASP09, WS96]. Adaptive [ABD^{+}12, AHMT08, BSS00, BKN07, BSG03, BSPB07, CJ11, CdLG^{+}09, DJW04, DGG^{+}10, EEM^{+}11, MGH09, NKPH08, NBC^{+}06, PBSB09, SF11, TNGK08, WAS06, ZBH12, DN04, LRS03, dBBC02, AKK03, BDIR07a, BDIR07b, Bou11, BDN08, BSS00, DN04, DJNR06, DLN10, DGSN11, DKN^{+}11, DLŠ11a, DLŠ11b, GG98, KTB09, KKA05, MMBND07, NKPH08, ND04, TT10, TNGK08, VKG12, WAS06, dBBC02, RSL01]. adaptivity [RTB02]. ADB [LR94]. ADB-94 [LR94]. ADBIS [ACJ08, BDG04, CE01, CIT11, EHKP05, EBT11, GMV09b, GKMN10, INR07, KMTW03, LMV98, MN02a, MPS06, MHW12, ŠPTM00]. ADBIS-DASFAA [ŠPTM00]. ADBIS’99 [ERW99]. ADC [AM91a]. ADCONS [TPCB12]. Addressing [Ste09b]. Adelaide [BB90, EL02b, Li11a, SCB01]. Adelphi [EFM04]. ADG [BR07a, HW06b, SNRG11, SZ11, GWY99, RGW01, Win04]. ADHOC [CSRS08, FLR11, KO07a, KR06, LPR12, NW10, RGLA09, SC05, NBK04, PBK03]. ADHOC-NOW [CSRS08, FLR11, KO07a, KR06, LPR12, NW10, RGLA09, SC05, NBK04, PBK03]. ADI [MTH08, MHD09, MDH10c]. ADL [Ada96]. ADM [GLS05, GZSC07]. ADMA [AGL^{+}07, CFZ10, CZF10, HYP^{+}09, LWD05, LZhL06, TLZ^{+}08, TKCW11a, TKCW11b]. ADMI [CGL^{+}09, CBG^{+}10, CBS^{+}12]. Administration [Sch74]. ADNTIIC [CFVV^{+}12, FdCLJ^{+}11]. ADS [ADMH04]. ADT [BRT11, RT09]. Advance [Bra80, SF10]. Advanced [ABK^{+}09, ABDW10, Alt01, BG03a, BMK08, BDPS02, BH07, BDZ08, BPZ09, BI11, BCP12, BIK01, BB00a, BJK^{+}05, CLLD09, CL04, CJ06, CEFZM05, CDCJL08, DO89a, DO89b, DFPS06, EEK^{+}87, EKLM87, EK12, Enc80, EFR08, FHJ^{+}02, Fil91, GT03, GLW07, GTVC06, GBMS11, GNTF02, GS08, GI03, HKP08, ILR95, IMW04, IMK06, JM95, KILW10a, KILW10b, KS04b, KKMN07, KHI06, KHS07, KFM07, KS00, LJS08, LJR^{+}10, LK99, LLWL04, LTW06, gLPZ^{+}12a, gLPZ^{+}12b, LM03, MM01, MSS03b, MP93, Nie94, NS98, PKS01, Pro11, ŠPTM00, WW07, YMY^{+}10, YKU11a, YKU11b, YYH^{+}12, ZYDL09, Zil87, AV99, ASP85, And95, BE74, Bau75, BE76, BSG84, BI03, BB05b, BJ86b, BP05, BRR87a, BRR87b, Bre97, CE95, CJ05, CB06, CJ95, CS01a, FHM82, FM96, GLIC10, GSB78, Gun94, HL04]. advanced [KSS02, KU94, KBD87, LSDP81, LPSD83, LMŠT01, MSH85, NM97, NY93, Nos88, PT97, RCT05, Ste94, TV87, VY93, WTZ^{+}04, ZOM05, AB93, AGL^{+}07, ABS91, AS98, BL08, BCG^{+}11, BTPPS05, BTPPS06, BTPPS07, BTBP^{+}08, BTPPS09, BTBP^{+}10a, BTBP^{+}10b, BTPP^{+}12, BvLR04, CNX05, CFZ10, CZF10, Cha02a, CDW86, CMV96, CP02, DYCD09, DHtM^{+}12, DGN01, DRS05, DKRT06, DGJ09, DP06, EM03, FS03, Foo99, GL92, GAB02, GNTF02, HHL07b, HHL07a, HWLJ08b, HWLJ08a, HYP^{+}09, HZBF10, HZGZ10, HGBF12, HGGG12, IHX^{+}08, JO99, JM95, JJ03, JDAW12, KPS09, KOS07, LMS96, LWD05, LZhL06, Lou92, MST92, MS03c, MR11, Mül10, NK99, Nos88, OP97, OYY03, PF05, PT98, Per10b, PS04b, PMWO02, RFBW12, RUL04, RRU05, RBC93, Sat97, SLL^{+}06, SSB90, SHO99, TLZ^{+}08, TKCW11a, TKCW11b, TYZ11]. Advanced [TPCB12, TBG05, VU05, WB00, WBW94, XZCL07, YLLZ04, ZJXC03, vEGW09]. Advances [AM91a, AT02a, ABG09a, AFR^{+}04, ABFR07, BE11, BM05a, BRR87a, BL04b, CSCO00, CT03, CHD07, CD00b, CK06c, CDND11, DMW00, DBB^{+}08a, DBHS10, DKP00, DLO03, EJPR01, ERRW03, FRS08, FPFR05, FKS07, GTP12, GF03, GdATM05, GRG06, GM07a, GM08a, GLM^{+}08, GK12, Gol90, GKN03, HBE07, HCF06, JKL02, JSK10, JVA04, KGRD10, KJB09, KKDV02, KKD04, KLC07, LO11, LKD08, Li11a, LKT08, LLO08, LK04, MHA09, MTWD06, MAFSS04, NL09, OFR09, Odl87, OT07, PRK12, Pom88, RWD09, SH12a, ShK06, SSAP05a, SSAP05b, SBM11, ŠPTM00, SCB01, TBD12, VKJ03, WR11, WY04, WVD02, WZ08, YWG04a, YWG04b, ZK06, ZT04, Ada96, ADR01, AD98, BBE01, Ber98, BDR99, BF00, BFP04, BGR03, BRR87b, BHN97, BDW97, BCC99, CCL02, Coe02, CYR02, CD01, DRR04, DKD05, DLM99]. advances [Esp01, Fra00, Fur05, GW96, GH96, HG98, LM00, LMW01, Len97, MS02a, NDF94, O’S03b, Ohl93, RR98b, RR98c, RBW95, SA97, VZ06, ZJ05, AP09, Abe10, ARSB09, ANS04a, ANS04b, ANS04c, AFK90, AD06, ARC^{+}07, AFM01, ABMH08, ACR07, AC11, Ami98, APS06, ABG96, AK12, ACJ08, AKR^{+}09, BYdVZ^{+}12, BTT90, BCD^{+}03, B^{+}12a, BS11d, BS11e, BL04a, BCM08, BBKP06, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BBP^{+}12a, BBP^{+}12b, BHGP02, Bel00, BCC^{+}10, BDG04, Ber08, BLB^{+}03, BSTL07, BCP^{+}04, BCI85, Bih03, BR02, BC85, BTKP^{+}11, BP01, BTZ^{+}10, Bon03, BC08, BK96, BMYZ95, BBMSD09, Boy01, BH05, Bra90, Bri93, BB02, BL11, CRJ11a, CRJ11b, CC04, CLKL09, CHKL10, CC05, CFB^{+}05, CE01, CVLL12]. Advances [CIT11, CL07, Cer07, CSM10, CCD^{+}06a, CCD^{+}06b, CL06a, CWC^{+}07, CP88, Che99a, CLY02, CCH02, CJ08b, CLZ^{+}09, CHG^{+}07, CWL01, CPC07, CM06, CFVV^{+}12, CFG^{+}11, CS02a, CAB10, Cop95, dRCVT11, Cra05, CGvR02, CHR10, DSZ04, Dam91, Dat09, Dav91, De 07a, De 95, DMB^{+}11, DFK91, Des94, Dig04, DCMS09, Dit88, DTW03, DLW^{+}07, Dwo06, ENP02, EKS10, ERW99, EHKP05, EBT11, EH95, Emb98, FKP^{+}05, FWY05, FK10, FMC04, Fei91, FGB11, FMS10, FIAP01, FdCLJ^{+}11, FMN99, FCLL02, Fra04, FLMS05, FLMK06, Fum97, FG04b, Fur95, GBH11, GJ09, GRT02, GPAD08, GKT09, Gil10, GSYZ10, GK95, GSV05, GMV09b, GKMN10, GW93, GQ95, GPP09, Gün88a, GS91, GHK^{+}10, GPL99, Had03, HRK^{+}07, Hal09, Ham00, HY12, HKB99, HKM95, HS00, Hel94, HM03, HSP^{+}12, HP09, HCL05, HK05c, HK05b]. Advances [Ho12a, Ho12b, Hof01, HhKRT06, HKT12, HO98, HZH05, HXC^{+}08, HCS11a, HCS11b, HSMPA09, IRM93, ISS^{+}06, INR07, IAL^{+}07, IK12, IN11, JOD98, JBJ94, JM02, Jen01, JJP^{+}02, JE06, JWbG^{+}06, JWG^{+}06, Jou09, KMTW03, Kal97, Kam99, KKS11a, KKS11b, KLZ07a, KCYL08, KL05, KS01, Kil01, KM96, KGV^{+}09, hKA10, Knu02, Kob96, KW07b, KKC09a, KKC09b, KI12, KKMN07, Kow01, KS00, Kra98, KBC^{+}09, KMS10, Kur07, LH11, LC06, Lai03, LMR^{+}06, LOX99, LM06a, LSW01, LSW02, LLCL05, Lau00, Lee04, LL07, LW11, LTL^{+}11a, LTL^{+}11b, LRG04, LWM^{+}08, LLLL08, LPC^{+}11, LLT^{+}04, LWF04, LST^{+}05, LZS^{+}08, LFP^{+}09, LCB12, LXH^{+}12, LJ00, LMV98, LCB97, LSL04, LLL06, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, LABM^{+}12, LRH10, LFL05, LGM11, LSW^{+}10, MOP^{+}08, Mah04, MSP^{+}09, MN02a, MEK03, MPS06, MTCOK10, Mat09]. Advances [MF08, Mau96, McC96, MZ98, MT04b, MEB05, MRSV07, MH04, MSS03a, MV91, Men07, MSW02, MN98, MR07a, MKR07, MNLM06, MS00c, Mor95, MK96, MHW12, MSX10, MWK^{+}09, MB97, Nao07, NZS^{+}06, NSS^{+}07, NKLC06, NRR12, NPS06, NR08, Nyb98, OiDP98, OS07, Oka00, OS03a, OA93, OF15, PS02a, PCH^{+}06, PZK07, PCA^{+}09, PB07, Pat11, PDMFF12, PpWR96, Per02a, Per05, Per06, Per07a, PS07a, Per08, PS08a, Per09a, Per10a, Per11a, Per12a, PBGK03, PJM^{+}08, Pfi01, PTM^{+}11, Pic86, PSN95, Pie08, PDG92, PJ12, PLK^{+}12, Pre00, PM09, QLXL07, QLK^{+}10, QLK^{+}11, QV90, Rab10, RM02, Rea01, RYG11, RBsC^{+}06, Rog11, RDKT01, RBGG06, Roy05, RRG85, Roz86, Roz87, Roz88, Roz90, Roz91, RRP91, Roz92, Roz93, Rue93, SNC12b, SW07, SGPP06, Sar03, SNE08, SSRA98]. Advances [SCM88, SMH^{+}12, SV97, Seb03, SP90, SZ93, SVA05, SOS96, Sho05, SLC01, SU97, SCR06, SAG10a, SAG10b, SMK01, SHP^{+}05, Sma97, Sma08, SF96, SC98, SCZ10, SPC^{+}08, SLKL09, SXC05, SO99, Ste99, Sti94, SM01, SZT^{+}08a, SZT^{+}08b, SKN05, SHŚ^{+}07, SSZ12, TM09, TSG00, TST10a, TST10b, TSCW11a, TSCW11b, TCHB12b, TCHB12a, TSJ12a, TSJ12b, TG04, TKCH09, TN09, Tor93, TGAH11, TDK^{+}10, Vau06, VMBMN04, WHL09, Wag08, WC95, WYL01, WCO05a, WCO05b, WCO05c, WLY05a, WLY05b, WLY05c, WYZ06a, WYZ^{+}06b, WS12, WYP12a, WYP12b, WSTI08, Wha03, Wie99, Wil86, WYX08, WD00, XCd03, Yak00, Yak02, Yak04, YN06, YT99, YLWY06, YX04, YSR^{+}06, YKL06, YHZ09a, YHZ09b, YHZ09c, Yun02, ZYRP10a, ZYRP10b, Zan00, ZdC08, ZJ06, ZSM^{+}09, ZLK10a, ZLK10b, ZHLW12, Zhe02, ZJL06, ZNCF03, ZLY07b, ZW09, ZYRH06]. Advances [dO98, dSK12, dSTP11, vdHHsHD06, vdHS10, vdHP12]. Advancing [EKS00, KLLP12]. ADVIS [YN06, Yak00, Yak02, Yak04]. advocating [MO97]. AE [Fon00, HLR^{+}98, LMCH12, All96]. AE-CAI [LMCH12]. Aegean [Mak86, Rei88]. AEGS [BBDS12]. AEMAS [MŠKL02]. AEPP [CPGL05]. AES [DRS05]. Affect [PB08a]. Affective [ADMH04, DGSM11a, DGSM11b, Pai00, PPP07, TTP05]. Affordance [RHD08]. Affordance-Based [RHD08]. AFP [KPS09, VU05, JJ03, SHO99]. AFPAC [SK97, SZ00]. Africa [ABGL09, BL10a, CP11, LW12, MV12b, NP11b, Pre09, Vau08, WW03, YJKSW10]. AFRICACRYPT [BL10a, MV12b, NP11b, Pre09, Vau08].

AFSS [PS02a]. Agadir [EML^{+}12, KS10c]. against [AKTT12, KT11]. Age [DTW03, DLW^{+}07, FWY05, IJ06, JMK09, LWF04, MSW02, MSS04, MdSC09, YKL06, WYL01, BGG^{+}12, CTYG10, GLW^{+}12, SPÖ^{+}10, WLO^{+}11, WJL^{+}12, LZ00]. Agent [AT07, APN08, BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, BPD06, Bv97, BDDS05, BDDS06, BDDS07, BGJ11, BBT10, BHH08, BLVV07, CW94a, CM95, CGG^{+}07, CDN12, CvRVSN12, DLFP02, DDE04, DSLT08, DSRW08, DSLT10, DSHL11, DS09a, DFPV11, DHtM^{+}12, DBB12a, DLW12, DP10a, Dig04, DvEF06, DFN10, DLGJ10, FBSMS08, FST09, FvdTDG12, GLZ^{+}03, GCL^{+}06, GB99, GGS09, HNH^{+}09, HENR03, Han02, HRT^{+}03, HRR^{+}06, HPS09, HMBD09, Hug03, IST07, IGN05, JSS08, JNHJ10a, JNHJ10b, JKN^{+}12, KM00, KL02, LB03, LOST04, LT05, LOTY05, LTÅ^{+}11, LMP04, LGR^{+}04, MMP03, MMV04, MVC07, MSZ09, MVL11, MPR07, MRPM10, MRP11, MPR12, MB09, Nag01, NGHJ07, NJHJ08, NS99, NVSB^{+}07, ONC^{+}11, OSV11, PSP^{+}02, PAV^{+}10, PMMR06, PM96, RPR08, RM09, RWSS12, SS08a, SSVW12, Sch01b, SS06b, SA06, SPON08, TGH06]. Agent [TT06b, THR03, TtHVS06, TNGK08, VP96, VSMS12, VASP09, WPM06, WPM07, WBD08, WJ95, YYI^{+}09, ZZ04b, AKK03, BK05a, BJ01c, CL01, CRW97, CGLR05, CS12c, Dam97, DM01b, DLT05, DG00, DC01, DL04b, DKN02, Had96, HWP01, HRTR05, JL00, KMTU03, KKA05, LMŠT01, MPB05, MT02b, MSY00, MWJ97, MSR99, MSS05b, Oss99, Pad99, PWZ98, PVP05, RMR05, Ras01, SCG98, SBD03, SRW98, Vei03, WR01, WS96, Wei97, WPM05, WMT96, vEHD05, BGHS05, BLVV07, CW01, CKTW12, CS12c, DS01b, DC01, DvEF06, FPRAW04, FRA05a, FS07a, GGS09, GMO03, GHSW04, GOW03, GGS11, HNH^{+}09, HR99b, JNHJ10a, JNHJ10b, JKN^{+}12, KBHSW06, KHSM^{+}08, KMTU03, LSJ06, LP08, LGS09, LS09, MSY00, MZ06a, NGHJ07, NJHJ08, NS99, ONC^{+}11, OGM05, PSP^{+}02, PZ07a, SS06b, WG11, WWC02]. Agent [ZZ04b, vEHD05, vEDA04, Bv97, DLT05, GB99, LMŠT01, MŠKL02, MMP03, PVP05, SCG98, ZL97, ZL98]. Agent-Based [HRT^{+}03, HRR^{+}06, LMP04, THR03, ZZ04b, DC01, HRTR05, MSY00, Pad99, Ras01, SCG98, Vei03, ZZ04b]. Agent-Mediated [NS99, PSP^{+}02, FPRAW04, FS07a, LSJ06, vEDA04, DC01, FRA05a, PSP^{+}02]. Agent-Oriented [BGHS05, GMO03, GHSW04, GGS11, KBHSW06, KHSM^{+}08, LP08, LGS09, MZ06a, PZ07a, WG11, CW01, GOW03, HR99b, OGM05, WWC02]. agent-related [KMTU03]. Agent-technologies [KMTU03]. AgentLink [DS01b, KBEP03]. Agents [ADP09, ABB^{+}10a, AOSV08, APV09, Bee02, BPD06, Bv97, BHH08, CBG^{+}10, CBS^{+}12, CW94a, CM95, DGR06, GB99, GGS09, GOZ05, GZSC07, GYA^{+}06, GDG11, Har11b, Hor00, HMG03, HKM^{+}07, Ish99b, KMD02, KjHGG11, KOS02, KOOL03, KOK04, KRP06, KHPS07, KPP08, KHK^{+}08, KVMH09, KL02, LSJ06, LB03, LwCM00, LY01, LS09, MB09, Nag01, NNPW12, Nar01a, OROD07, OPT01, OPP04, PWZ98, PPCS12, PMA^{+}07, PTZ03, Pic01, PLI08, RABR03, RKNV09, RME04, SS06b, Sur02, TT10, TNGK08, VP96, Vig98, VKMT11, VKG12, ZS00, AG98, Alb99, AKK03, BK05a, Bau00, BHGP02, CL01, DAB01, FBSMS05, GLS05, JL00, KK97, KW98, KSW99, KK00, KZ01, KBEP03, KM00, KKA05, MS99a, MT02b, Mul96, MWJ97, MSR99, NcC99, NA97, OTZ01, PGA^{+}05, PM96, PG01, RPZ97, RH98]. agents [SRW98, WR01, WJ95, WMT96, YY01, BBDS12, BDG^{+}12b, CGL^{+}09, CBG^{+}10, CBS^{+}12, DJS06, DBSvD09, Dig11, JDMB08, JDMB10, KI01, KjHGG11, MK03b, MSS05a, MBA05, Neh99, NRW04, PWZ98, Paz03, VP96, CN98, Oss99, TP97, von92]. AGI [STL11]. Agia [MT05a, MTH05]. Agile [AMS06, EB04, MW03, WW02, BMH05, MS03a, ZEL04, CDSS07]. Agility [MNW08, WPR07]. Aging [LBM^{+}10]. AGTIVE [SNZ08, SVV12, PNB04]. AGTIVE’99 [NSM00]. Aguascalientes [GM07a]. AH [DN04, HMPZ09, NKPH08, WAS06, BSS00, RTB02, dBBC02]. AH-3 [RTB02]. Ahead [MSWZ06, BMJ88, Wil01b]. AI [AB11, APS06, BE11, BM05b, BS06a, BJP12, Ber08, BFP04, BL11, DVVA08, DBB^{+}08a, DBHS10, ELLP12, FK10, FKS07, Fur05, GJ09, GPAD08, GK12, HBE07, KL05, Kim05b, KW07b, KPK^{+}10, KI12, KMS10, LM06a, LMT10, MPV12, MHA09, PDMFF12, PW92a, PT08b, SCR06, SS06e, VW07, vE91, vHH09, BBE01, BB90, BF00, CPSA10, Coe98, CS02a, ETG02, FD93, GRT02, GF03, GKN03, Ham00, JKL02, LRG04, Len97, Li11a, LZ02, MS02a, MS00c, NL09, OT07, PPCS12, RR04a, ShK06, SA97, SM01, SCB01, TG04, VS02, VP04, WR11, WY04, WZ08, XCd03, ZJ05]. AI*IA [BM05a, BP07, CT03, Esp01, LM00, PS11a, SC09a, AGS91, GS95, Tor93]. AI09 [GDG11]. AI’93 [Yao95]. AI’94 [Yao95]. AI’96 [McC96]. AI’97 [PWZ98, Sat97]. AI’98 [AS98, MN98]. AI’99 [Foo99]. AIASTS [NSM07]. AICI [DWWL09, DMLW11a, DMLW11b, DMLW11c, LWDM12, WDGL10a, WDGL10b]. AICOL [CPSA10, PPCS12]. AICOL-I [CPSA10]. AICOL-I/IVR-XXIV [CPSA10]. AICOL-II [CPSA10]. AICOL-II/JURIX [CPSA10]. AICOL-III [PPCS12]. AICS [CF10, OSR^{+}02]. Aided [AO02, AP04b, BJ06, BM09a, BL02b, CBF01, DH07a, DPA07, GW98, GQ11, GM08b, HJ00, KP15, MDY^{+}10, MS12a, MDP04, MDPQA09, MDPQA12a, MDPQA12b, TCJ10, AH96, BHH74, Cou93a, DPA05, DF96, Dil94, ES00a, Enc80, ER05, Gru97, HP99a, HV98, HM04, HS03, LS92, MDBF01, OK96, PMD90, PMD92, PMD94, PMDA96, PMD97, PMDK00, SC96, Wol95, vP93, CK91, SLF91]. AIE [AE05, AD06, CHCA09, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, JDAW12, MMO^{+}11a, MMO^{+}11b, NBGA08, OA07, CHA03b, HA02, LPA00, MVA01, OYA04]. AIE-92 [BR92]. AIE-98 [MdPA98, dPMA98]. AIE-99 [Ima99]. AIED [BBKM11]. Aiello [SS06e]. AII [AJ94, Jan89, Jan92]. AIMDM’99 [HSL^{+}99]. AIME [BAHH07, CSAH09, MHK05, PLC11, Ria08, RtTMP10, RtTM12, BSW95a, DKB03, Ker97, QBA01]. AIMS [BB07, CCBW11, HS08b, SP09, SNČ^{+}12a, SD10]. AIMSA [DD10, DPT08, ED06b, RA12, CD00a, Sco02]. AIMSA’98 [Giu98]. AIN’97 [MSRP98]. AINTEC [CJ05, CJ06, FS07b]. Aires [AMGJ12, CN05, FC04, OdQ12]. AIRS [CKLN10, HNS^{+}12, LYMM05, LSL^{+}09, LLM^{+}08, MZWZ05, NLKJ06, SSO^{+}11]. Airth [ZS05b]. AIS [GLS05, GZSC07, KKGK11, Kim05b]. AIS-ADM [GLS05, GZSC07]. AISA [Cha02a]. AISB [Fog94, Fog96, CS97, Fog95]. AISC [ACR^{+}08, ACD^{+}10, BC04b, CIW06, JCC^{+}12, CBC^{+}02, CRL01]. AISC’98 [CP98a]. AISMC [CC93, CC95, CCP96]. AISMC-1 [CC93]. AISMC-2 [CC95]. AISMC-3 [CCP96]. Aix [Oli07, BTQ02, HCT02, KWA02, TL02, Vää99]. Aix-en-Provence [BTQ02, HCT02, KWA02, TL02, Vää99]. Aizu [Bha00, Bha02, Bha05, Bha07, BB03, HRA02, KSB10, KMSB11, OS03a, WVDS05, YGGJ04, GY03]. Aizu-Wakamatsu [Bha05, Bha07, KSB10, KMSB11, OS03a, WVDS05, YGGJ04, GY03]. AKA [BMS08b, BJM09, Sch06]. Akademgorodok [BBP96, BBZ00, BBZ01, BZ03]. Akko [EK81b]. Akron [CGBZ08]. al [CW94a, WPD04]. ALA [TT10, VKG12].

Alagoas [LVP04]. Alain [PQR11]. ALAMAS [TNGK08]. Albacete [COB06]. Alberta [JPBT05, JRSN09, KH02, LOH05]. Albufeira [BPP91]. Alcalá [MRU12]. ALEA [NSM07]. Alegre [GPP09, PBK^{+}10, dO98]. Alejandro [SPG10]. ALENEX [BS01c, GM99, MS02c]. Alessandria [BM10]. Alexandria [CCJS07, Jaf99]. ALFA [BBPP09]. Algebra [AM91b, Arn90, BMS04a, BMS08b, BJM09, BS01b, BL07b, BAH09, Buc85, Cav85, DN88, Dav89, DO89a, DHMM96, Ehr85b, EEK^{+}87, FILP06, FZ86, GMV06, GMV07, GMV09a, GKMV10, GKMV11, GKMV12, HS02, LOS05, MWD06, Rao92, Sch06, AB81, AP83, BMP90, BC88, Cal82, CMM93, CGM95, DG01, Fos99, FHP03, GMV05, Hee94, HP89, Jan88, Kir96, MMR91, Mor89a, MM97, Pol86, Sak91, Tis94, Zip92, vH83, CKC11, FHRR05, FJM06, KLT09, MMH07]. Algebraic [BMS04a, BM03b, BS01b, BL07b, BAH09, CR01a, CM09c, FS07c, Fit84, FILP06, KG12, Kuh10, MK12, Mos04, Pal09, WPH03, dS11a, BCM00, CMM93, CGM95, Dav81, Fia99, FMO05, Fos99, FHP03, Gia89, HP89, MMR91, Mor89a, MM97, MO97, Ng79, Par98, Sak91, ZWS81, ACG^{+}07, AN95, AHK07, BCG02, BMP90, BF91, BKL^{+}91, BR07b, BR09, Bre91, Cal86, Coh92, Coh94, GLW88, Gue81, Hae99, HR96, HHM97, HRRY08, HNP12, Joh97, JV06, JP11a, KW90, KL92, KR02b, KR11, LA94, MR08, RMS04, Rus00, SK97, SZ00, Win11, WB89, WN96]. Alghero [TBG05, TN09]. ALGOL [Bra76]. Algorithm [AF06, AI98, FMS02, GK90, Gud08, HK04, Hal00, Kap10, KL88, KL96, MHS10, NU92, PS02b, SS94, AF06, AI98, BFMS01, BS01c, FK12b, GM99, Gud08, HK04, Hal00, Kap10, KL96, MS02c, MHS10, NW01, NU92, PS02b, SS94, VZ99]. Algorithmic [AJ94, ALZ11, BDFM12, Erl06, EHMM09, JP07a, LOH05, SLSW12, EK77, Fri93, Haj96, AKZ01, AH94, AJ94, AJS00, BLS06, BDCM04, Bos00a, BRT11, B^{+}12b, Bue04, Buh98, Che10, CP06b, Coh96, Dol09, Dos93, Fek08, FK02, FX08, FGTZ08, GJT03, GLZZ09, GKS^{+}07, GZ09, HMT10b, HPP06, HST07, HSVZ10, Jan93, JL95, JSZ95, KL07a, KSUZ11, KKS10, KCK08, LM97, MP09, MXZ05, MS08b, NR04, NR06, Per11c, Ric98, RT09, Ser12, SA96, SJT05, WY99, van97b, vdPS08]. Algorithmics [ALZ11, FMS02, LCY10, PF07a, PWY08, SLSW12, DHX09, Pol86, LWZ09]. Algorithms [AvLR91a, AL11a, AFvZ02, AAN09, AS03, AJRS03, BS09, BJ10, BDL08, BDIR07a, BDIR07b, Ber12, BK02, BG10b, BS01b, BL07b, Bra07, BAH09, CJRR07, CPZ07, CLZ06, CEQ^{+}11, CCM08, CPP07, DU09, Dem07, Die04, DLŠ11a, DLŠ11b, DDS02, EP06, EK07, ENO12, EK81a, EM10, ER12, FZJE^{+}05, FZJE^{+}06, Fes10, FKM09, FK12a, FILP06, GJRR08, GJRR11, GH07b, GJRS12, HHJ^{+}86, HP06, HHLR08, HGZ05, IS11a, IS11b, JK00, JLV02, JSO04, JKRR04, JSO11, JP04a, KS08a, KK11b, Kla12, KTB09, KP09, KS00, KKL12, LHDD08, LBDW09, MSS11, McG08, iNR08, PKM^{+}10, PR11, Pat00, RF10, RiN12, RR08, RM04a, RK09, SKS^{+}06, Sch10a, SSJR10, SOP12, Ste01, STWS07, SBH07, SBH09, SG01, TZS00, Vah09, WZJX12, WZ09, WMB10b, WVDS05, AJM87]. algorithms [AS06, BM96, BJK06, BR89, BN91a, BM94, Bri95, Bri97, But91, Cal86, CJRT05, Cle99, CMM93, CGM95, DG96, DJRZ06, DJNR09, Dji89, FS03, FR95b, Fer96, FW98, Fos99, FHP03, Fur95, GH03, Ger04, GJRT01, GS00b, HZ05, HR95, HJRS99, Hof82, HKNW07, HL91, HP89, JR98, Jen03, KS02c, Koj91, LS89, Lom93, LKZCS05, Mak86, MS96a, MY01, MMR91, MT97, MO92a, Mor89a, MM97, Nik05, NS81, Oli00, PS04a, PR87, PSO05, Rei88, Sak91, Sch93b, SZ92, Ste98, Tei93, TV94, Tha84, TC98, TSK92, vL88, vLS91, AR99, A^{+}08, Akl95, AR04, AHW07, Asa90, Asa96, Asa06, AiNOW11, ADL09, AE06, BP03, Bil98, BC07, BJ12a, BBD97, BGP00, BCP94, BM02b, BZ08a, BL05c, BM06, BW97, Cai08, CFI06, CD10, CCP10a, CCP10b]. Algorithms [CI98, CL08b, DSS89, DSS91, Deh93, Deh97, Deh99, DST01, DSS03, DLOS05, DSZ07, DGST09, DIS11, DH11, DD06, DZ03, DS96, DDI09, DZ94, ET01a, EMO10, EF12, ENO12, FS09a, FT04a, FHP11, GM01, GH07a, GG02b, HM08, HNF08, HYPY10a, HYPY10b, HC09, Iba92, IKO03, JRPP07, JK04, KL95, KS10b, LT00, Len93, Leo04, LIJ97, MPS98, Mel04, MSS03b, Mey01, MR02, MS10e, MC05, Nes99, Ng93, PPS03, PS11b, RT12, SW09a, Sch10a, Spi95, SWB06, Sta95, The89, Tok07, WW07, XCHZ11b, XCHZ11a, XSA^{+}12a, XSA^{+}12b, ZZ02, dBM10a, dBM10b, vL94a, Bau00, FP90, PS93]. Algorithms-ESA [Mey01]. ALGOSENSORS [Dol09, ENO12, Fek08, KCK08, NR06, Sch10a, NR04]. Alicante [CO94, EV07, FIAP01, GTVC06, MSA99a, MSA99b, MMM05, MMM11, PCG^{+}07, VMBMN04]. Alive [BKT12]. Allan [Koh06b]. Allocation [SWB06]. Alloy [DFG^{+}12, FGK^{+}10]. Almaty [VV03]. ALP [HR96, HHM97, LA94]. ALP’98 [PGM98]. Alphen [HM11b]. ALPS [LJS08]. ALSIP [OBS12]. ALT [AJ94, BLS06, B^{+}12b, Dos93, FGTZ08, GLZZ09, HST07, HSVZ10, KSUZ11, SJT05, AKZ01, AJS00, BDCM04, CBNR02, GJT03, Jan93, JSZ95, LM97, Ric98, SA96]. ALT’99 [WY99]. Alternating [Wal82]. Alto [GW98, NW97, SW09b, SC96, ZZ04a, ZG03, dKIM^{+}07]. Altrincham [BBP89]. AM [TYNM05, Dit88, PPCS12]. Amagi [MB82]. Amalfi [BC05, CGP03, GD10, ODV08, Per11c, Tuc01, VD12]. AMAS [VMvN12]. AMAS-BT [VMvN12]. AMAST [BR97, JV06, JP11a, KR02b, MR08, RMS04, AN95, Joh97, Kat99, Rus00, WN96]. AMAST’98 [Hae99]. Amazonia [Hae99].

Ambient [APP12, CSPC09, Jac09a,

LMMN10, ORB^{+}09, SPCG07, SP07, Ste07b, ACvLdR03, ACdR^{+}08, BHV11, Cai05, CA06, KMS^{+}11, MEAC04, MKB^{+}09, PdRM^{+}12, SDG^{+}07, SS05b, TdRM^{+}09, dRWK^{+}10]. ambiguity [Fra96]. Ambleside [TN03]. AMBN [OBM11]. AMCP [NK99]. AMDO [PF06, PF08, PF10, PFM12, NL01, PH02, PD04]. AMEC [FRA05a, FS07a, LSJ06, FPRAW04]. Amendment [TDB^{+}06]. America [AB10, HN12, OPS11]. American [BYGP95, BBR07, Coe98, CHK06, FC04, FB12, GPAD08, GPV00, KMS10, LBNF08, LRG04, LO10, LM98c, MSAdAS05, MS00c, PDMFF12, Raj02, SCR06, Sim92, dLWC03, GRT02]. Americas [FGH98, FT04b, Ric02, Whi00]. Ames [BMP90, HMW^{+}08, HS98b]. AMET [NS99]. AMET-98 [NS99]. AMFG [ZGT05, ZZTG07]. Amherst [JP12b]. AmI [ACdR^{+}08, KMS^{+}11, PdRM^{+}12, SDG^{+}07, TdRM^{+}09, dRWK^{+}10]. Amir [MP10a]. AMITA [NSM07]. AMKM [vEDA04]. AMPLE [DHtM^{+}12]. AMR [BDN08, DJNR06, DLN10, DGSN11, DKN^{+}11, MMBND07, ND04]. AMSTA [HNH^{+}09, JNHJ10a, JNHJ10b, JKN^{+}12, NGHJ07, NJHJ08, ONC^{+}11]. Amsterdam [ABD^{+}09, AFvZ02, ABdJS10, AKR^{+}12, BK90, BG91a, BBdB12, BSC12, BAHH07, BCT06, BLvdH10, BL06c, BAWH00, CCJ07, CP88, CW99, CA10, Cle99, CLT05, DvV96, EBSS98, EK10, Feh11, Fin99, FGK^{+}11, GH09, Gue04, HKB99, HZ10, Hee94, HW01, HHW01, HB11, HS99, Jäh99, KvETvdH10b, KvETvdH10a, KMS^{+}11, Knu02, MV06, MBH^{+}11, Ria08, RKNV09, SL11, Sam76, Sch04a, SBH98, Slo99, STDH02c, STDH02a, STDH02b, SCH04c, SHP^{+}05, SHS10, Swi99, Tho99, Vad07, WJ95, dBBGdR06, dBBGdR07, dBBGdR08, vEGW09, vE91, vL88, vL89, vdBK05]. AMT [ALPH10, LWYN09, ZCGH11, LYhL^{+}01].

AMTA [FT04b, Ric02, Whi00]. AMTA’98 [FGH98]. Anacarpi [BGR08]. Analaysis [BBA12]. Analogical [AJ94, Sch03b, Vel94a, Jan87, Jan89, Jan92]. Analogy [Neh99]. Analyses [FZJE^{+}05, FZJE^{+}06]. Analysing [Han81]. Analysis [AL11a, AG09, AJRB09, ARSB09, ABB^{+}11, ABF^{+}07, AGM05, AG07, ABG09b, AG11, AV08, AMGJ12, AMPT09, AL11b, AB06, BCE09, BS06b, BLB^{+}03, BSTL07, BG03b, BC10b, BC10c, Bra07, BH00, BBH08, BH11, CK04a, CK04b, CK06b, CK06a, CK08, CK10a, CK10b, CK12b, CK12c, CJ08a, CCK^{+}08, CH10, CZZ08, CAB10, CGS02, CM11, DJAP07, DV09, DFP^{+}12, DIP12, DPSU11, DPNM12, DFGN12, Ekl04, EP07, EM10, EV09, ER12, FT11, FR09, FK12a, FG01, FG04a, FSV09, GBH11, GSW05, Geu98, GSYZ10, GR06b, GZ06, GH07b, GI03, HP06, HP02b, HK11, HKT12, IS12c, JP04a, JP09a, JN12, KPK05, KC07, KC09, KC11a, KC11b, KŽ04, KP09, KH11c, KKH07, KS07c, KS10c, LN04a, LWZ09, LPZ12, LR09, LSIT11, LSW12, LHK02, MDHB11, MCKB06]. Analysis [MF11a, MF11b, MVW09, MdlBP05a, MdlBP05b, MBMS07a, MBMS07b, MTOK06, MO08, MS12c, MS06b, MNLM06, MSS02, MO06, NYHO07, NZS^{+}06, NSS^{+}07, NF07, OMMSP06, OV09, PM08, PS09a, POP09, PvES12, PCPS03, PBH03, PS07a, PS08a, PSD12, PW03, PY05b, PBCF08, Pri03b, PP04, RR08, RT07, RDPMA^{+}11a, RDPMA^{+}11b, RSB07, REF^{+}06, RMB10, REPH06, RMK07, RSK08, SHJ09, SK11, SRS03, SH12b, SLZH07, SW10, SKK04, SRM^{+}11, TWYhL01, TCYZ12, VJ11, VZM^{+}10, VSH11, Wan04, WB09, WPZA11, Yah11, YLS^{+}12, Yi06, ZSM^{+}09, ZJL06, vES10, AW05, BMSMS04, BBH^{+}05, BGH06a, Bee89, BBH01, BDV95, BE05, Bri95, Bri97, BHK01, BS06c, Can89, CM12, Chi95, Cle99, CF99, Cou93b, CS96, Cou01, Cou03, CRYS09, CMM01, CK93, Del97a, Del97b, Den85, FS03, FKP^{+}05, GP05]. analysis [GG05b, GH03, Gia04, GS00b, HZ05, HKB99, HS05a, HS95c, Hof01, IS98, KC05, KS02c, KHG01, LY04, Le 94, wLN99, Lev98, LVW05, LCB97, Mac06, MS96a, MY01, MD04, MS00a, MB97, Myc95, Nak92, Pal00b, PT05, PVC^{+}99, RV05, STC04, SL05, Ska01, SL99b, SDP97, SG96, Ste98, Uhl82, Van97a, VWY97, VWY01, And95, AHSC11, Dam97, DHM00, EVV^{+}11, Ful75, Kon04, ZGT05, ZZTG07, Eve75]. Analysis/Synthesis [ZJL06]. Analytic [BM11b, EF02, GW09, MP03b, Oli07, BHP95, Bec05, Dyc00, Gal97b, MMMO96, Mur99, de 98]. Analytical [Rei99, BLW82, ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12]. Analytics [CFH^{+}12, SB12b, SBM08]. Analyzing [SBH07, SBH09]. Anaphor [Bra07, DBM09, HDBM11]. Anaphora [Bra07, DBM09, HDBM11, Bra07, DBM09, HDBM11, Hir81]. anatomy [Sac04]. ANB [HNP12]. Andhra [PSDS11a, PSDS11b]. Andratx [PCPS03]. André [BCM08]. Andrei [BBP96, BBZ00, BBZ01, BZ03]. Andrews [BW05a, CHD98, Coh10, Dyc94, Dyc00, OWM04]. Andros [Kut98]. Angeles [DLW06, EJM10, EG97, GGHS07, HMP08, PS09a]. Angers [HLC^{+}12]. Angra [MEB05, RM04a, SW07]. Anguilla [DR06, Fra01, Fra99, Hir97, Hir98]. Animals [AHMT08, DGG^{+}10, NBC^{+}06, Tu99, ZBH12]. Animation [ERK07, LZ02, Tu99, LZ02]. Animats [AHMT08, DGG^{+}10, NBC^{+}06, ZBH12]. ANIREM [BPD06]. Ann [SS11a, SW75, BWT95]. ANNA [Luc87]. Annecy [AD06]. Anniversary [AM03, BV92, LIBP07, BZ83, Van89]. annotated [BKL^{+}91]. annotating [Luc87, SKP07]. Annotation [FKM08, MMM10, MF06, GF12b]. ANNPR [MST12, PMS08, SM06, SE10]. Annual [ACL12, AF02, AH03, AP10, AH97a, AT02b, AGI^{+}06, AHW07, Atl08, AM05, AE06, ABFR07, BM03a, BE11, BS11a, BA07, Bat09, Bel00, Ber10, BFP04, Bon03, Bra02, BVNW87, Bra73, BHN97, Bri93, BG07b, Bun96, BCC99, CL06b, CC09a, CL90, Cho12, CS00b, Cop95, CW88, Cra05, CBCGA12, DL06a, DV10, DH11, DBB^{+}08a, DS96, DH04, DBHS10, DH07b, DT06b, Dwo06, EFW93, EMW94, EF12, Ési06, EIRW06, FRS08, FC98, FL08, FJ92, FR99a, FJ10, Fra04, FKS07, Fri01, FD11, Fur05, GKL09, GM11a, Gil10, GK12, GH96, GGS99, GK09, GAM^{+}11, GV09, GMV12, GKN03, Hal09, HW01, HSP^{+}12, HBE07, HG98, HM96, HW08a, IZ02, JRSN09, JW05, JKL02, Jou09, KM08a, KS12a, Kil01, KS02d, KLFK10, Kra98, KU09, LO11, LV06, Li11b, Lin07]. Annual [LS06c, MZ07a, MT04a, MZ04, MP95a, Meh85, Men07, MHA09, MMK^{+}05, MVN86, MC89, Müh75, Nao07, NDF94, Ngo09, NT98, OFR09, OT11, Ong05, OF15, Pat00, Pat11, PJ12, PF07a, PWY08, PR96, Rab10, RT00, RM97b, Rog11, RBW95, SNC12b, SW03a, STS04, Sho05, Sie75, Sma08, SH07, Spi95, Sti94, SBH^{+}04, TS10, TW07, VY01, VW08, Wag08, WL04, Wie99, Yun02, dBM10a, dBM10b, vB97, AR04, AL01, Apo92, Apo93, ACP05, Asa99, BYCC03, Bil98, BL05c, BW97, yCW96, CM04a, CG94, Des94, DZ03, DD05, DL95b, Du00, FR01b, FS09a, GU95, GS00a, HM08, HA00, HK98b, JL97, Kal97, Kob96, Len93, MT99a, Mey01, MR02, MMK98, Nes99, NH03, PC99, PN93, SMD04, ST01, TM99, Wan01, Wan05b, WZ03, vL94a]. Anonymity [Fed01]. Answering [AMCL02, AYB^{+}09, CHAL04, CDY^{+}11, EK12, LPOA^{+}06, Nau02, ACL98]. Ant [DDS02, DBB^{+}04, DGB^{+}06, DBB^{+}08b]. Antalya [ADK04, YŞ03]. Antibes [BP04a, Fau90, AF02]. Antibes-Juan [AF02]. Anticipating [Bee00]. Anticipation [PBCF08]. Anticipatory [BSG03, BSPB07, PBSB09]. Antipolis [BF08, Ber00, BJN03, CSRS08, CDT11, DL06a, Dit01, FLS06, FZJ01, HW03, HSMPA09, HDLM09, JVA04, KMP84, MMM01, SRG04, dBBM09]. Antonio [DH94, Goe98, Gup99, MO92b, PS97a, RA00]. ANTS [DGB^{+}06, DBB^{+}08b, DBD^{+}10b, DBB^{+}12b, HMT10b, HPP06, vdPS08, DBB^{+}04, AH94, Bos00a, Bue04, Buh98, Coh96, DDS02, FK02]. ANTS-I [AH94]. ANTS-II [Coh96]. ANTS-III [Buh98]. ANTS-IV [Bos00a]. ANTS-IX [HMT10b]. ANTS-V [FK02]. ANTS-VI [Bue04]. ANTS-VII [HPP06]. ANTS-VIII [vdPS08]. Antwerp [BTPPS05, BTPPS06, DGM08a, DGM08b, KR10b, Par84, Tri98]. AOIS [ALS^{+}05, BGHS05, KBHSW06, KHSM^{+}08, GHSW04, JP03]. AOM [Bru06b]. AOSE [DHtM^{+}12, GGS11, LP08, LGS09, MZ06a, OGM05, PZ07a, WG11, CW01, GMO03, GOW03, WWC02]. Aotomata [UMN^{+}08]. AP2PC [BDG^{+}12b, DJS06, JDMB08, JDMB10, MSS05a, MBA05, MK03b]. APCHI [LCHS08, MJR04]. Apizaco [GRG06]. APLAS [Hu09, Kob06, Ram08, Sha07, Ued10, Yan11b, Yi05, Chi04, Oho03]. APNOMS [AH07, HTMC09, KT06, MCA08]. applicability [She03]. Applicable [BC88]. Application [Abe10, ACH00, BPV04, BCI85, BMCCM11, Boy01, BGH06b, CP88, CHZ08, CEER96, Dam91, Dav91, De 95, DS11c, DW06, DW07b, Fum97, GH12a, GQ95, Gün88a, HM11b, Hei12, Hel94, HŽ07, HZB07, IY06, IRM93, Jac09b, JP03, Kur07, LC06, Lai03, LOX99, Lee04, LW11, Man09, Mat09, Mau96, MS11a, MR12c, OiDP98, Oka00, Pfi01, PSN95, Pie08, Pre00, Qui04, QV90, Roy05, Rue93, SHØ06, SC11, SZ93, TO08, WS12, WR09, Zhe02, Bra76, CM03a, CER79, DOSY05, ENR83, Ehr87, EKR91, EEKR00, FLS06, FKS86, GLIC10, GHMT05, ID03, KS95, MD01, Nyb98, Pic86, Ste99, WTZ^{+}04, WW03, YP01, AB97, BR96a, BDR99, DD95, DS98, EL02b, HP12, Jen92, KD99, Kno98, Mar93, NS00, Val94]. Applications [AK10, AM10, AFvZ02, ADDL08, ACL12, AVP07, ABKT11, ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12, AS03, AGL^{+}07, AE05, AD06, AMGJ12, ABJ02, AGS^{+}11b, AGS^{+}11a, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, AAR05, Atl08, AJN12, Baa07, BW05a, BS11c, Bar08a, BA07, Bar03, BGP11, BCE09, BHHK08, BWAM09, BMS04a, BMS08b, BJM09, BCV06, BJSW09, BKW08, BKW09a, BG06, Bih03, BG12b, BGS^{+}03, BMSW12, BPT06, BC10b, BK02, BCGY09, BGPR10, BIMW12, BKM01, BDGN07, Bra07, BKW06, BHLvdP07, BHQ10a, BHQ10b, BZ08b, BLVV07, BDFS04, CC04, CCL06, CG04, CL11, CFZ10, CZF10, CP10a, CD00a, CZ05, CY03, Cha02a, CHCM11, Cha05b, CH04, CLLD09, CC09a, CWZ11b, CLZ06, CEQ^{+}11, CCdS03, CHCA09, CEFZM05, CCAS12, CHA03b, CSY09, CY11b, CS12a, CCM08, Cra05, CBCGA12, Cur09, DLFP02, DL06a]. Applications [DYCD09, DVVA08, DSS11, DIMV09, DFMV10, DMV12, DV09, DBM09, DPP09, DD10, DVV07, DDK11, DMB^{+}07, DPT08, DFPS06, DT06a, DP97, DXZ07, DHP09, DFW^{+}11, DKOZ05b, DT07, EDD04, EV01, EM06, EK10, EMO10, ED06b, FPP11, FR11, FKB02, FSdlPT11b, FJ10, FKBP11, FKM07, FMRW09, GTT04, GSW05, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, GG07a, GG07b, GG07c, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, GMP06, GN12b, Gie05, Gil10, GT04, GPE05, GHSY06, GH12b, GHP09, GNTF02, GHH06b, GS08, GH93, GV09, GAR04, GRR05, GYD^{+}06, GKOT00, HNH^{+}09, HRK^{+}07, HLSZ11a, HLSZ11b, HKP08, HG00, HK05a, HJ12, HDBM11, HA02, HRSS08, HKLP12, HM05a, HRM^{+}10, HMMW10, Hor00, HMG03, HHL07b, HHL07a, HWLJ08b]. Applications [HWLJ08a, HYP^{+}09, HJL^{+}09a, HJL^{+}09b, HZBF10, HZGZ10, HGGG12, HGPH12, HMJG12, IR08, IKR00, Ima99, IR07, IHX^{+}08, Jac07a, Jac11d, JNHJ10a, JNHJ10b, JNH11a, JNH11b, JLS11a, JLS11b, JKN^{+}12, JDAW12, Jou09, JY12, KSS08, KMD02, KGL04, KLM^{+}07, KM96, KYL07, KILW10a, KILW10b, KY07b, Knu02, KKM^{+}11, KPK^{+}10, KFMM06, KKMN07, KLFK10, KSGT10, KR04b, KV11, Kul09, LGK^{+}04c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, Lau12, LLWL04, LTW06, LYY07, gLPZ^{+}12a, gLPZ^{+}12b, LY12, LOS05, LZhL06, LYT06, LFIM07, LHDD08, Li11b, LSTZ12a, LSTZ12b, Lin12, LSK08, LBDW09, LDXZ10a, LDXZ10b, LC11, LMC12, LPA00, MOP11, MOH^{+}12, MWD06, MKPV05, MPV12, MP06, MSG07, MG08, MZ07b, MSZ08, MNZ09, MS06a, MS10b, MS10c, MS12b, MVW09, MRŠ03, MMV04, MPB05, MVC07, MSZ09, MVL11, MSS07, MVOF08, MP10b, MTOK06, MMP07]. Applications [MMO^{+}11a, MMO^{+}11b, MT08c, MLTC11, MLK04, MCSN10, MM04, Mid01, MSA99b, MP01b, MÁ05a, MFÁ^{+}09a, MP12, MOT12, MVA01, MMM05, MC98, MPS04, MMM11, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, Nac12, Nao07, NP12b, NGHJ07, NJHJ08, NBGA08, NHJ12a, NHJ12b, ONC^{+}11, OT11, OA07, OW09, Ölv10, Ong11, OYA04, OF15, OSRS07, PHR09, PKS01, PCN10a, PCN10b, PCN10c, PKM^{+}10, PB07, PAC08, Pat11, Per02a, Per07a, Per08, Per09a, Per10a, Per11a, Per12a, PNPP12, Pfe06, PJ12, PPH07, Pro11, QSBC98, RCB^{+}04, RLL04, RS09a, RA12, RW06, RK09, Roc07, RMK07, RSK08, SKS^{+}06, SS11a, SK11, SH08, SKP10, Sch06, SSN08, Sco02, SHGB05, SG10, SO09, SWJX12, Shu11b, SM11b, SCP^{+}12, SMEGA10, Sma08, SPO11, SDH03].

Applications [Ste01, Ste07c, Ste09c, Ste11c, SWB03, STY^{+}07, SKM07, SS10, ŠPTM00, SRL12b, SD06, Tah00, TWYhL01, TLZ^{+}08, TKCW11a, TKCW11b, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, TGH06, Tis02, Tre09, TDK^{+}10, UWB^{+}06, VTK^{+}12, Vau06, Vor08, Vos03, VP04, WRP07, WZD11, WZJX12, WZH^{+}12, WZ09, WL04, WMB10a, WCBD08, WD10a, WD10b, XXX^{+}09, XHXL08, XLHX08, XYZU11, YNB08, YDW08, YSL12, YHV12, YMY^{+}10, YY09, YLLZ04, YKU11a, YKU11b, YYH^{+}12, ZS00, ZCDT10, ZYDL09, ZZLL10, dSOSR04, dSOSR06, vO04, Abr01, ABL96, AMU03, AG98, Alb99, AM91c, ADW05, BSMM04, Bac00, Bai05, BJK06, BR92, BCST99, Ber99, BG93b, BN91a, BBP93, Bla80, Boo91, BM94, Bra80, BRR87a, BW01, Cag00, yCCZ07, CYGL04, CO94, CC91, CPGL05, Chi95, CFG96, CG88, CW89, Com97, CMT01, CN98, DH97, DIMV11, DJ90]. applications [Der89, DCP95, Dim03, DLW06, Dur12, Ell95, FBSMS05, FL00, FM97, FJL97, Gae96, Gan96, Gir99, GZR02, GPPW08, GN90, GAR03, GS94, GY03, HI93, HT97, HCT02, HWP01, HRR93, HS94, HG96, HK98a, HSSY05, HPR05, Hof03b, HKNW07, Hsi95, HL00, HS98d, JW05, JSdSA05, Jou85b, JS94, Kar94, KKM04, KT96, Kir93, KM00, KMTU03, KKGT03c, KKGT03a, KKGT03b, KW96, KA05, LRS03, LSW00, LG99, Les87, LWD05, LVW05, LSS^{+}04, LY05, LMŠT01, LCG97, LKZCS05, LIH99, MSS12, MŠKL02, MMP03, MMT97, MKT98, May01, MXZ05, MC94, MdPA98, Mit98, ML95, NR99, Nie03, Nip98, Oli00, PWZ98, PS04a, PCG^{+}05a, PS98a, Per05, Per06, PM96, PVP05, PG01, Pin89, Pla96, QS00, Reu97, Reu99, Reu01, RT95, RB00, Sal83, STC04, SL05]. applications [SWAF95, SLL^{+}06, SL99a, SKY06, SW05, SRL12a, SRL12c, Tah01, TLC00, Urz05, VDPD96, VS02, VWY97, VWY01, WT96, YK82, YY01, ZL97, ZL98, ZLJ05, ZZO03, ZOM05, dBdRR93, dPMA98, AS11a, ADM^{+}12, Ber02a, Ber12, BHVV09, Boe01, BF04, Cag02, CZZ08, CCNY10, Cho02b, CGG^{+}07, CD05, CBS02, CK01, CR04b, DFKM08, DCC^{+}10, DBD^{+}10a, DCC^{+}11, DBD^{+}11, DAC^{+}12, FW09, FBHW00, GLZ^{+}03, GCL^{+}06, Gia07, GBC^{+}08, GBC^{+}09, HKW^{+}11b, HKW^{+}12b, HP98, JMC10, KMPH09, KMS12, KP11, LBSBW98, LP10a, LR94, LGR^{+}04, MMP07, MS10d, MZF94, NSM00, Pau01, PB08a, PNB04, Rai03, RR98c, RBC^{+}05, RBC^{+}06, SNZ08, SVV12, SHGB05, SHW09, SG01, WS06, vHV08, vdAB03, AY96, AAN00, BCBG05, BF91, BD93, BSA05, BKS96, CGLR05, Gre94, Jah93, MLW93, Nol88, NA97, PR87, PD96]. applications [Ral81, Rei94, Sch03c, WB89]. Applied [AE05, AD06, CHB^{+}12, CWiM^{+}09, CHCA09, DCMS09, GO96, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GL74a, GL74b, HA02, HSTU99, JDAW12, MAD^{+}04, MMO^{+}11a, MMO^{+}11b, NBGA08, OA07, OYA04, PR98a, TNA^{+}10, WDPW02, WDPW04, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, dPMA98, BV92, Bli87, CHA03b, Man75, APFV09, BSZ12, BDPS02, BGKY08, BS01b, BL07b, BAH09, CMM93, CGM95, DMW96, DMW06, FHJ^{+}02, Fos99, FHP03, FILP06, HP89, HSTU99, IKY05, JYZ04, Jón12a, Jón12b, Kåg98, KEDW07, KY07a, LM96, LT11b, MAD^{+}04, MMR91, Mor89a, MM97, Pol86, Sak91, SMMG01, Waś96, ZYH03, ZYB06, ZY10]. Applying [NJA^{+}11, NMP^{+}04]. Approach [FHW98, Geu98, MÁ05a, Sto04, Ull08, Aso04, EKT94, God96, Jac91, Koj91, Lou94, Mat91, MO97, Naz94, Paz97, PT92, Rus05, Tha84, Vei03, Bar02, BHS07, KSF^{+}03, Kuh10, BBEdK98, Hof03a, Mul96, Nag96, ST96, Sip97, Wal99b]. Approaches [BH06b, BBLT06, BS06b, CPSA10, Cer05, CW09, DL07, FI08, GO11, GW06b, HNRW06, HRT^{+}03, IMW04, IMK06, Jac11a, LMT00, Liu11, MMO^{+}11a, MMO^{+}11b, Mel07, MJ07, NT06, PPCS12, PU04, RK09, RT10, SKRZ04, SC09b, Urs02, WMS04, Yu02, dLZ12, BB01a, CKM93, CRYS09, DA01, Est98, Fin99, FK95b, God05b, GN05, HRTR05, HP99c, Hus01, Ima99, KS91, KBD87, LPA00, Mai00, NZ03, Pez03, Ras01, RS96, TvdBI02, WK02, BDMS08, CN06, NcC99, SKP10]. APPROX [CJRR07, CJRT05, DJRZ06, DJNR09, GJRR08, GJRR11, GJRS12, SSJR10, AJRS03, GJRT01, HJRS99, JR98, JK00, JLV02, JKRR04]. Approximation [AJRS03, CJRT05, DJRZ06, GJRT01, JK00, JKRR04, KH02, Len04, LRS98, MPS98, MR07b, BJK06, CW00, HJRS99, Rol97, RV02, Sud95, ZS05b, AJRS03, BS09, BJ10, CJRR07, CJRT05, DJRZ06, DJNR09, EP06, EK07, GJRR08, GJRT01, GJRR11, GJRS12, JR98, JLV02, JSO04, JKRR04, JSO11, KS08a, PSO05, SSJR10, SOP12]. approximations [Klo94]. APPSEM [BDPS02]. APPT [CNX05, DGJ09, TYZ11, XZCL07, ZJXC03]. April [AL11a, Abr07, AM91b, AM91a, AZ96, ADDL08, AM07, ADB^{+}05, Ama08a, ACR07, ADL05, Ang97, Apo92, AGI^{+}06, AT02c, AS07a, AS10, AS12, Arm02, AKR03, Auc98, Aya95, BCG02, BGHJ07, BYGP95, BYdVZ^{+}12, BBP89, Ban98, BLW09, BP04a, Bar11, BH08, BDIR07a, BDIR07b, BBB07b, BCD04, BdR94, Ber10, BMS08b, BB05b, Ber98, Bet83, BCI85, BB89, BJ86a, BD97b, Bir12, BHL90, BSBM05, BHR90, BKV10, BHHJ11, Bod91, Bod05, Boe01, BDS04, BI04, Boo91, BL02a, BBMSD09, BHT97, BLV02, Bra93, BHH^{+}06, BBH08, Bri97, Bro94, BV05b, Buc85, BLS02, BC96b, BC96c, Cag00, Cag02, CSCO00, CJ11, Cal82, CV04a, CZ02, CVTP12, Cav85, CR01a, Cer05, CP88, CWY^{+}06, CDLZ06, CPZ07, CJ08b, CMS08a, CLZ^{+}09, CLLD09, CYCL09, CMM09b, CWL01, Cho12, CCLR98, Chr99]. April [Chr00a, Chr01, CCMR02, CCMR03, CCMR05, CCMR06, CCMR07, CCMR10, CMMR11, CMSA12, CH96, CW99, CPPB06, CV10, CU08, CFG^{+}11, Coe02, CSTD08, CTE^{+}06, CS97, CvH07, CC09b, CM10, Cro03, DKP05, Dal92, Dav91, DH97, DS05b, De 07b, DRDKE11, DIMV09, Deg03, DG12, Dei90, DBPZ05, Der89, DCM82, DCP95, DCC^{+}10, DBD^{+}10a, DCC^{+}11, DBD^{+}11, DAC^{+}12, DPT12, DR81, DS03, DLŠ11a, DLŠ11b, DFPS06, DPSU11, Dov05, Dro08, DFW^{+}11, Due04, DL07, Dyc94, EOE^{+}07, EM08, EFG^{+}09, Ehr91, EKR10, EAES^{+}10, EHMM09, Est98, FC04, Fau90, FZJE^{+}05, FZJE^{+}06, FR95a, FR96, FR97, FB12, FM04, Fia99, FI08, Fie02, Fis09, Fit93, FK12a, Fog94, Fog95, Fog96, Fon03, FM84, For88, Fos02, FS10, Fri94]. April [FTF89, Gac02, GH03, GH08, GJ93, GH94a, GH94b, GW03, GBC^{+}09, GVB12, GO11, Gie05, Gir99, GLIC10, GPV00, GP12, Gor03, GR06a, GS00b, GM04b, GH07b, GAN05, Gue90, Gyi96, HM94, HW06a, HZ05, Han98, HM12a, HLKX12, Hed03, HNM02, Hen08, HS98a, HP06, HAB98, HS97, HRTR05, Hof11, HM01, Hor02, Hsi95, HM98, Hus01, HU05, Ibr95, IHX^{+}08, JP05a, JLM00, JP04a, JPD04, Jou09, Joy03, KG74, KS04a, KS02c, KOL^{+}04, KTC^{+}05, KSW05, Kir96, KILW10a, KILW10b, Kno11, Knu02, KTB09, Kos98, KKMN07, KYT01, KP10b, LBNF08, LMR^{+}06, LW95, Le 02, LTW06, gLPZ^{+}12a, gLPZ^{+}12b, LLP^{+}07, LYT06, LFP^{+}09, LCHS96, LYCV08, LG95, Lom97, LST^{+}12, LO10, LD05b, LD05a, LM98c, MOP^{+}08, MRC12, MMMP12, Mai00, Mai88, Mai90, MT09c, MNS05, MP03a]. April [Mal08, MSS11, MMR07, MY01, Mat02, MT04b, Mel86, MH11, Mic90, Mil00, Mil01, Mio90, Mit05, MP12, MS08b, MAFSS04, MTU09, MSK^{+}12, MBS05b, Mos01, Mue01, Nac01, NZ06, NR98, NKLC06, NKJ11a, NKJ11b, NE02, Nie96, Nip98, Niv98, O’B12, OW07, OF15, PAD^{+}06, PBK^{+}10, PS12a, PR84, Pez03, Pic86, PMD92, PSNS99, PRG09, PRG10, PRG11, PKSS02, PJ12, Pol00, PZH96, QCDM06, QV90, Rai03, RCB^{+}04, RG05, Raj09, Raj02, RR08, Ras01, Reu97, RMB10, RSL01, Rob74, Rob80, Rol98, Rol99, RPZ97, RBC^{+}06, RF99b, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, RSK03, RSW12b, Sad13, Sag05, STP^{+}10, San01, San94, Sas05, Sch95, STT11, SKP07, SUW02, Sch04b, Sch01a, SPZ90, SE10, Seb03, Sei07, Sei12, SYY^{+}08, SWJX12, SFN^{+}11, Sim92, SBH98]. April [Slo99, STDH02c, STDH02a, STDH02b, Sma08, Smo00, SH07, Ste98, Ste95, SG07, SC00, TDWG11, TLC00, TKCH09, TNA^{+}10, Tis94, Tiu00, TNVDF06, TKHS05, UPB07, Urz05, Val06, VGM^{+}09, VMB^{+}12, VW08, VS02, WS02, WK08, Wal04b, WZH^{+}12, Wat00, WK02, WMS04, Wha03, Wil01a, WL93, WKK98, XYZU11, YZC^{+}07, YGE12, YMY^{+}10, YLLZ04, YKU11a, YKU11b, YYH^{+}12, YDKM06, ZTY^{+}05, ZYBX08, ZG03, ZZ99, ZZO03, ZOM05, ZYDL09, Zui99, dKIM^{+}07, dLZ12, vDKS11, vW97]. APWeb [CWC^{+}07, CLZ^{+}09, DLW^{+}07, DFW^{+}11, IHX^{+}08, LFP^{+}09, SLL^{+}06, SWJX12, WZH^{+}12, ZTY^{+}05, ZYBX08, ZLS^{+}06, ZZO03, YLLZ04]. APWeb/WAIM [CWC^{+}07, CLZ^{+}09, LFP^{+}09]. Aquila [CF09b]. Arab [SSO^{+}11]. Arabic [DJ08]. Aranjuez [Oqu07]. Arbor [SS11a, SW75]. ARC [BWAM09, BCV06, CCAS12, DMB^{+}07, KKM^{+}11, SMEGA10, WCBD08]. Arcachon [Tau08]. Architecting [dLGR03, dLGR04, CdLG10, Gie10, RSS06, dLGR05, dLGR07, dLDG^{+}08, dLFG^{+}09]. architectural [PS97a]. Architecture [AVP07, AAR05, BG10a, BL05b, CPC07, CGB11, DW06, DW07b, GO06, HK05a, MO05b, MBF08, MSUB04, Nar01a, OS03a, OWM04, Oqu07, PHR09, RW06, SH08, SUW02, Smi06, SXC05, VMvN12, YX04, FL00, GH93, HR99a, Hug91, JE06, Jou85a, Kah87, LSDP81, LPSD83, PS98a, RR89, SL99a, SKB98, Str99, YI91, ZL96, BMSHW09, BFBS11, BUHS08, GSW06, HRB12, LTT07, MSKY10, von79]. Architectures [BPR08, BWAM09, BDW08, BI03, BCV06, BDSW09, BZ08a, aCF06, CCAS12, CNmWH^{+}05, CKLM09, DKS^{+}07, DMB^{+}07, GPS03, GKS05, HPTV05, HKP10, HGZ05, HCR06, HYPY10a, HYPY10b, HC09, JRPP07, KKM^{+}11, KS10a, KS12c, MGH09, ORS89a, ORS89b, OSRS07, PFD^{+}10, PPH07, SEO^{+}09, SMEGA10, Ste07a, SDK^{+}08, Ste09a, Ste11a, Ste11b, VWH06, VBH07, WAW01, WMdLA10, WCBD08, WJ95, XCHZ11b, XCHZ11a, XSA^{+}12a, XSA^{+}12b, dBNT87a, dBNT87b, dMS06, AvLR91a, AvLR91b, AMU03, AJM87, BL94, BD93, BRW93, CL01, Coi99, DTV02, ES92, GSP01, Gut94, Hal94, HS94, JL00, Mak86, MT02b, MMR93, MWJ97, MSR99, PV04, Rei88, RMS^{+}05, SRW98, Ste94, TAS05, WMT96, YM01, vdL98, vdL01, BKB03, MGH09, MP95b, ON80]. ARCS [BL05b, BMSHW09, BFBS11, BUHS08, GSW06, HRB12, LTT07, MSKY10, BK80, MSUB04, SUW02]. Ardenne [FS07c]. area [BB89, MSH85]. Areas [AKS09, BY07, BGS11, HH04, JRSN09, MZ04, MV12a, PT06a, VY01, AMW07, HA00, JY14, NH03, ST01, TM99]. Arequipa [FB12]. ARES [QBY^{+}12, TQYX11, vdL98]. Argeles [Cho88]. Argeles-village [Cho88]. Argentina [AMGJ12, CFVV^{+}12, CFVCF^{+}12, CN05, FC04, FdCLJ^{+}11, HOC07, KMS10, OdQ12]. Arges [PPJRN^{+}10, Pau03]. ArgMAS [MPR07, MRPM10, MRP11, MPR12, PMMR06, RMR05, RPR08, RM09]. Argonne [LO88]. Argumentation [CPSA10, MOT12, PPCS12, MPR07, MRPM10, MRP11, MPR12, PMMR06, RMR05, RPR08, RM09]. arguments [van90]. Arithmetic [CS07, HH10, ÖRH12, vzGIK08]. Arizona [Apo92, BG93a, CTT93]. Arles [MP93].

Arlington [EKT94]. Armenia [DV07b, GKMV10]. Armidale [Yao95]. ARMS [DHtM^{+}12]. ARPA [Heb95]. array [FKS86]. Arrays [WLS89, GH93]. ARSPA [AL11b, DV09]. ARSPA-WITS [AL11b, DV09]. Art [DMW06, DRSS01, KEDW07, MRC12, PR98a, ALM95, BFS92, MPS93, PGV93, BKB03, FW98]. Art-QoS [BKB03]. ARTDB [AH98]. ARTDB-97 [AH98]. ArtDeco [ABD^{+}12]. Articulated [NL01, PH02, PD04, PF06, PF08, PF10, PFM12]. artifical [FD93, MS95, CML^{+}10, HNPP07, RST^{+}10a]. Artificial [ABG09a, AL04, AE05, AD06, ARC^{+}07, AFS02, AIL12, APS06, AP11, AMNT06, AN06, BV02, BE11, BM05a, BCD^{+}03, BPP91, BSW95a, BB90, BP07, BS11d, BS11e, BAHH07, Ber08, BR02, BV12, BCC07, BJ01c, BL11, CPS05, CSPC09, CRJ11a, CRJ11b, Cai11, CT03, CVV08, CV10, CSAH09, CUdlC04, CAP08, CWO^{+}09, CRS10, CKW11a, CKW11b, CSA^{+}12a, CSA^{+}12b, dRCVT11, DBB^{+}08a, DV07b, DBHS10, DKB03, EL07, ETG02, FST01, FK10, FV10, FSdlPT11a, FSdlPT11b, Fil91, FvdHKL06, FKS07, GJ09, GF03, GPAD08, GdATM05, GRG06, GM07a, GM08a, GK12, GM06a, GKN03, HWP01, HKI^{+}09, HA02, HV06, HBE07, HZ08, HLW08, HDGK11a, HDGK11b, HSL^{+}99, HHL07a, HWLJ08a, HJL^{+}09b, HZGZ10, HGGG12, IPSK04, Ima99, IS02, JN10, JKL02, JDAW12, KKS11a, KKS11b, Ker97, KW07b]. Artificial [KI12, KMS10, LM06a, LRG04, Li11a, LPA00, LLMR09, LGM11, LIBP07, MOBS06, MM89, MRSV07, MHA09, MMG10, MHK05, MCP93, MS95, MSA99a, MSA99b, MP01a, MP01b, MA03b, MA03c, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, MS00c, MVA01, MAFSS04, MSY00, NMM10, NSM07, NBGA08, NL09, NV01b, OBS12, OA07, ORB^{+}09, OBM11, OYA04, OT07, PWZ98, PCH^{+}06, PDMFF12, PW09, PLC11, PBCF08, PBSB09, PA03b, PS11a, QBA01, RAW07, RCS10, SHN07, SPCG07, SINK08, ShK06, SC09a, SCR06, SAG10a, SAG10b, SAUL11, SCB01, TG04, TNN^{+}01, TN04, TNM05, TNVDF06, TNY07, TN08, TNI09, TND10, TNYL11, VV03, VDÉ^{+}12b, VDÉ^{+}12a, VP04, WR11, WSN^{+}06, WSTI07, WY04, WZ08, YW06, ZdC08, ZL97, ZL98, ZGY04, ZO10, dCHM12, vdMS11, vM94, AZ96, APO96, AS98, AGS91, BBE01, BV05a]. artificial [BA99, B^{+}12a, BL04a, BR92, BCD05, BJ86b, BFP04, BHR90, BK96, BHN97, BCC99, Cad95, CSCO00, CC95, CFB^{+}05, CHA03b, Coe98, Coe02, CS02a, CC97b, DDFM05, DdCF98, Esp01, Fle02, FMN99, FG96, Foo99, Fur05, GRT02, Gom01, GS95, GH96, Ham00, HR91, HG98, Kar91, KL05, KJ91, KS01, KS93, Kow01, LM00, LM98b, Len97, MPP94, MPF95, MST92, McC96, MS02a, MN98, MMMD97, MdPA98, Mit98, MS00b, Mor95, NDF94, Nos88, Ohl93, OAdGBP00, Oss99, Pad99, Ral94, RM97a, RS99a, RBW95, Sat97, SM01, SV05, Tor93, VY93, VS02, WC95, Wei97, XCd03, ZL96, ZJ05, dO98, dPMA98, APPE09a, APPE09b, All96, ATO^{+}09, BSW95a, BAHH07, BLJ08, BC06, BBKM11, Bra09b, BWT95, BC04b, BF04, CC93, CCP96, CP98a, CBC^{+}02, CIW06, CRL01, CW94a]. Artificial [CD00a, CGK^{+}12, Col02, CSAH09, CF10, DVVA08, DWWL09, DMLW11a, DMLW11b, DMLW11c, DDI10a, DDI10b, DDI10c, DD10, DPT08, DKB03, DBH01, Dor02, DKOZ05b, DKOZ05a, ED06b, Fon00, GGHN97, Giu98, HLR^{+}98, HLC^{+}12, HMTH10, HDGK11a, HDGK11b, HSL^{+}99, JPBT05, Kay03, Ker97, Kim05b, KSDO06b, KSDO06a, KPK^{+}10, KRH09, KNK08a, KNK08b, LWDM12, LCF^{+}04, LNS11, MPV12, MST12, MSZ90, MHK05, MÁ05a, MTC^{+}08, NCBT04, OSR^{+}02, PLC11, PMS08, Pri91, QBA01, RA12, RSTZ04, RTZŻ06, RTZZ08, RST^{+}10b, RKS^{+}12a, RKS^{+}12b, Sav06, STL11, SM06, SE10, Sco02, Smi98, TLC06, Tan93, TBH03, Tu99, VDÉ^{+}12b, VDÉ^{+}12a, WDGL10a, WDGL10b, WV99, dCVK07, dSADM07a, dSADM07b, von96]. Artificielle [Col02, CML^{+}10, HLC^{+}12, LCF^{+}04, MTC^{+}08, TLC06]. Artiminio [DPNM12]. ARTIST [BS05]. artistic [HAB98]. Arto [KMPR04]. ARTS’97 [BR97]. ARTS’99 [Kat99]. ASA [KM00]. ASA/MA [KM00]. Asahikawa [HKI^{+}09]. ASCM [Kap08]. Ascoli [BMUV10]. ASCoMS [OD12b]. Ascona [Jen03, Nie94, Smi06, WNE99]. Asia [AH07, CWZ^{+}11a, CWYC12, CLY02, CYCL09, CChL^{+}10, CC12, CKLN10, CPC07, CKH10, DSZ04, DLW^{+}07, DFW^{+}11, HCL05, HTMC09, HNS^{+}12, HCS11a, HCS11b, JE06, KT06, LYMM05, LCHS08, LSL^{+}09, LLM^{+}08, MCA08, MJR04, MZWZ05, NKLC06, NLKJ06, OS03a, SSO^{+}11, SWJX12, SXC05, TCHB12b, TCHB12a, TKCH09, WSTI08, WKK98, XCR11, YZC^{+}07, YKF97, YX04, YLLZ04, ZYRP10a, ZYRP10b, ZTY^{+}05, ZYBX08, ZLS^{+}06, ZLY07b, CWL01, McK99, TLC00, Wha03, ZZ99, Zho01, ZZO03]. Asia-Pacific [AH07, CC12, CPC07, CKH10, DLW^{+}07, DFW^{+}11, HTMC09, JE06, KT06, LCHS08, MCA08, OS03a, SWJX12, SXC05, XCR11, YKF97, YX04, YLLZ04, ZTY^{+}05, ZYBX08, ZLS^{+}06, McK99, Zho01, ZZO03]. ASIACRYPT [Abe10, Kur07, LC06, Lee04, LW11, Mat09, Pie08, Roy05, WS12, Boy01, IRM93, KM96, Lai03, Oka00, PSN95, Zhe02]. ASIACRYPT’98 [OiDP98]. ASIACRYPT’99 [LOX99]. ASIAN [Cer07, Dat09, Mah04, OS07, ACN^{+}07, Bai05, BMC08, CCM^{+}04, CP98b, CP98c, Chi04, CJ05, CJ06, DA08, FS07b, FNPW05, GCSR07, GPYD09, GSV05, HS00, HO98, Hu09, JY96, JM02, KL95, Kap08, KKS11c, KKS11d, KKS11e, KKS11f, Kob06, Li96, LFK^{+}02, MAD^{+}04, MSG06, NNS06a, NNS06b, Oho03, PCN12a, PCN12b, PCN12c, Ram08, Sar03, SZC^{+}03, Sha07, SU97, SHRM06, Ued10, YKKZ07a, YKKZ07b, Yan11b, YT99, Yi05, ZiTM10a, ZiTM10b, ZiTM10c, ZW09, HS00, HO98, JY96, JM02, Sar03, SU97, GCSR07]. ASIAN’99 [YT99]. AsiaSim [Bai05]. Asilomar [CG94, GHR89]. ASM [BGR03, GKOT00, ZT04]. ASMTA [ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12]. Aspect [KOFJ09, KMK10, KMSJ11, LCH^{+}12, RA06b, RA07a, RA07b, RO09, RA06a]. Aspect-Oriented [KOFJ09, KMK10, KMSJ11, LCH^{+}12, RA06b, RA07a, RA07b, RO09, RA06a]. Aspects [ARS03, ACG^{+}07, ABDW10, AF02, AH03, AÖ12, ALZ11, BL12, BCC06, BDE12, BD12, BVNW87, CP10b, CDGW10, CP11, Che10, CP06b, CL90, CW88, Dai07, DGM09, DG10, DEG11, DH04, DMRS06, DMRS07, Dol09, DT06b, EDGP11, EFW93, EMW94, Erl06, Fek08, FR01b, FJ92, FHY08, FX08, FM84, GS09b, GZ09, Han07, Hen06, HHL07b, HHL07a, HWLJ08b, HWLJ08a, HJL^{+}09b, HZGZ10, HGGG12, HW08b, JP07a, Jay04, JLW07b, JMT11, KL07a, KR02a, KR04a, Kar09, KOFJ09, KKW11, KKW12, KR02c, KCK08, LM09, LOH05, MP95a, Meh85, MT99a, MVN86, MC89, MG07b, MMK98, NR04, NR06, PvES12, Per07a, Per08, Per09a, Per10a, Per11a, Per12a, PP11, PC00, PR96, RT00, RK06, RM97b, Rob11, RL11b, RD12b, RZ12, SLSW12, Squ09, TW07, VEG^{+}11, Yam08]. Aspects [vES10, AI97, BL05b, BSBM05, DW03a, Dam99b, DMR01, DD05, Ehr89, Gup99, HM94, HC05, HW05, IM91, Kar91, KSS02, KP01a, LLP99, LB90, LA05, Moo01, Ram01, Ret97, SS89, SW75, BBGR03, Dal92, DK90, JPR04, MM09a, vEV12]. Aspen [SWAF95]. Assembly [ÁCD^{+}12, DGG^{+}11b]. Assessment [BRS^{+}07, BL06d, FB09, HS07, HB11, KJ10, Zam08, JK05a, RBS93]. assets [GT05]. Assisi [LGK^{+}04c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d]. Assistant [NPW02]. Assistants [KKMW07]. Assisted [AJNJ12, AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, BHH04b, BHH04c, BHV11, DK02a, DK02b, DG05a, EP03a, EP03b, FMP11a, FMP11b, FMP11c, JNPV10a, JNPV10b, JNPV10c, LNS06a, LNS06b, LMCH12, MAFS08a, MAFS08b, NJ10, ORB^{+}09, TY11, TGM97, YHR^{+}09a, YHR^{+}09b, DDJ00, DG05b, Mau89, NV01a, NS90, TC99, Tom92, WC98]. Assistive [MKB^{+}09, Mit98]. Associated [GKMN10]. Association [AGS91, BM05a, BP07, BdBK03, BCC07, CT03, CUdlC04, DP10b, Esp01, FGH98, FT04b, GS95, LM00, Len97, LGM11, MOBS06, MCA^{+}07, MMG10, PS11a, Ric02, SC09a, Tor93, Whi00, ZZ02]. associative [FKS86, Hau86, Stu85]. Assurance [DLdS04, PCA^{+}09, GSP01]. AST [hKA10]. AST/UCMA/ISA/ACN [hKA10]. ASTD [CPGL05]. ASWAN [CWC^{+}07]. ASWC [ACN^{+}07, DA08, GPYD09, MSG06]. ASWSD [BKM06, BKM08b]. AT4DL [BCG^{+}11]. ATAL [MWJ97, CL01, MT02b, WMT96]. ATAL’97 [SRW98]. ATAL’98 [MSR99]. ATAL’99 [JL00]. ATC [CYM^{+}11, NRWI09, RJS^{+}08, XYM^{+}07, XBS^{+}10, YJMU06]. Athens [AOSV08, ACHH12, AKSO06, BDL08, CL11, Chr91, CDA^{+}06, CLBP11, Del00, EKF12, EM06, GANACL11, GPT10, GHMV11a, GHMV11b, GHMV11c, Hal94, HPP88, Kar94, KSDO06b, KSDO06a, KPK^{+}10, KKS10, KA05, LCDT08, MMMP12, MKR^{+}04, PS04a, SH12a, Sel95, SS06d, SZ05, TNP^{+}07, VEG^{+}11, XW11]. Atibaia [MS00c]. Atizapán [GM08a]. Atlanta [ABS01, ASE^{+}07, BBD^{+}10, FS10, FHP11, KMR^{+}05, MZ07b, MSZ08, Pic01, SR11, SvASD05a, SvASD05b, SvASD05c, WGL96, YD98, vdMS11]. Atlantic [FMC04, MSS03a, SKPG12, SKN05]. Atlases [CPR^{+}10, CKP^{+}12]. ATMOS [GKS^{+}07]. Attacks [BSC12]. Attention [Fri06, PR07, PT09, PTRH05, PR07, PT09]. Attribute [Uhl82, AM91c, DJL88, DJ90, Yel88]. ATVA [BC10c, BH11, CCK^{+}08, CM12, GZ06, LR09, NYHO07, PT05, Wan04]. Auckland [BMS^{+}08a, BK05a, BBKM11, CCD04, DKN12, HRK^{+}07, HZK08, JP08, KPS01, KŽ04, KKC09a, KKC09b, PSST07, SK08, WZ08, ZGY04]. Audio [AJB09, CKOB11, MSB12, MB06, NSFB10, OB07, PB08b, SBC^{+}94, Her92b, LG95, Ran93, BCB97, BS01a, KJR05, KN03, Ram98]. Audio- [BCB97, KJR05]. Audio-and [BS01a, KN03]. Auditory [YAKMJ10]. Augmented [CHCM11, DSL08, LEP^{+}10, SR07, SEG09, SF11, YJ04, YJS^{+}06, LMCH12]. Augsburg [MSUB04]. August [AH00, AD05, AB10, AMW07, ARSB09, AW01, Akl95, ACD^{+}07, Alb99, ADB^{+}12a, ADB^{+}12b, AGL^{+}07, ANF08, ABR09, AR10, AH96, Ama08b, ABD^{+}99, Ami98, ALPH10, ADW05, AFGvE10, AFGvE11, And95, ABK12, ATO^{+}09, AGT98, ABD08, AJRS03, ABW10, ACR^{+}08, AKS09, Baa03, BM03a, BG10a, BK90, BG91a, BM99a, BW03a, BH06a, BDK10a, BKH11, BJ06, Ban92, Ban93, Ban94, BGG^{+}12, BK05a, BUG93, BS76, BBP04, BCM08, BPV04, BMR^{+}04, BHRU09, Bel00, BCST99, BLJ08, BNUW09, BF12, BLB^{+}03, Bes93, BND01, BG12a, BKW09a, BC93, BG06, BK05b, BY07, BG12b, Bil98, BD98, BGS11, BSS11, BTPPS07, BTKP^{+}11, BBRS05, BGU94, BLKW01, BKLP11, BFMS08, BFJM09, Bon03, BF06, BS93, BS03b, BR07a, BFMR96a, BFMR96b, BFT^{+}08, BvLR04, BDGK00, BDGN07, BW01, BCH^{+}05, BMS01, Bri93]. August [BJKK02, BHLvdP07, BHQ10a, BHQ10b, BFMS01, BHH74, BSS00, BW92, BP93, Bru04, BCF^{+}02, BR96b, BC98, BE01, BD03, BT06, BR07c, BTF05, BFKO05, BFK^{+}08, CAD^{+}02, CCS04, CC93, CC95, CCF^{+}08, CMT02, CF00a, CM95, CHPW05, CRW97, CP11, CAS^{+}09b, CY03, Cha02a, CKK^{+}09, CJRR07, Cha02b, CPC^{+}99, CJRT05, CL06b, CLZ06, CEQ^{+}11, CFK^{+}12, CPC07, CY11b, CM04a, CJK88, CP99, Cle92, CS00b, CH98, CGK^{+}12, CKOB11, Cop95, CKC11, CF10, CvRVSN12, CBBS09, CDG89, CVÉ07, CM05b, CD11, DGT10a, DGT10b, DL06a, DLV04, DSHL11, DV10, DN04, DFPV11, DG09a, DSS89, DSS91, Deh93, Deh97, Deh99, DST01, DSS03, DLOS05, DSZ07, DGST09, DIS11, DS01a, Des94, DKN^{+}11, DJRZ06, Die03, DR02b, DMRS07, Dim03, DDK11, DLdS04, DJNR09, DLGJ10, DSL08, DHO08, DS11c, DGG^{+}10, DTW03]. August [DMW96, DBH01, Dor02, DGJ09, DW06, DXZ07, DER94, DZ94, DL95b, Dwo06, EJM10, EDGP11, EL07, EH95, EF02, EEM96, Ell95, EMT02a, EMT02b, Ern07, EN74, Esi93, Eta05, ET06, FPP11, yF75, Fer96, Fer98, FMS10, FIAP01, FKK04, FO95, FvdTDG12, FCLL02, FKW08, FG96, Fra04, FCD^{+}04, Fre01, FM99, Fre96, FMRW09, FD11, FG04b, FS06, FLP11, Fur95, GVO^{+}10, Gae96, GTT04, Gal97a, GM00, GWY99, GML^{+}08a, GLSY10, GLW^{+}12, GY04, GG09, GM01, GL10a, Gec81, GFV09, GG07a, GG07b, GG07c, GHP^{+}11, GPR^{+}12, Gia04, GM12, GSYZ10, GN88, God05a, GJRR08, GJN10, GJRT01, Goe98, GA09, GJRR11, GK02, GLM93, GLM97, GGS99, GYA^{+}06, GGK^{+}09a, GGK^{+}09b, GAM^{+}11, GH93, GC81, GRW86, GVT^{+}11, GMV12, GF97, GS91, GJRS12, GDG11, GPV88, Hal09, HLSZ11a, HLSZ11b]. August [HWW^{+}10, HKB99, HH04, HWP01, HAM95, HK98a, HG00, HK92, HPV00, HNRW06, HMP08, HM03, HSP^{+}12, HA00, HK10b, HSKB09, HJRS99, HhKRT06, HSB^{+}10, HRRY08, HNP12, HK98b, HSB^{+}96, HZH05, HLI06a, HLI06b, HLI06c, HHL07b, HHL07a, HYP^{+}09, HZBF10, HZGZ10, HGBF12, HGGG12, HGPH12, Hug91, HM05b, IZ02, IY06, IS98, IS02, JPBT05, JRSN09, JW05, JKRR04, JP07a, JSWhT11, JNR11a, JNR11b, JS05, JJ03, JL97, JP04b, JP05b, JP08, JP09b, JP12a, JP94, JY14, JS94, JQ04, JY12, KPS12, Kal97, KKP02, KC07, KR10a, KTH09, Kar83, KT96, KLP04, KLB^{+}06, KLS10, KBP07, Kil01, KYL07, KGV^{+}09, Kin08, KSW99, KOOL03, Kob96, KP99, KP01b, KS04b, KH02, KMSS02, KRRBV07, KBH03, KMS12, Kow01, KU06, KN09b, KP05, KT11, Kra98, KMS89, KMYJ08, KSH05]. August [KKH07, KK07, KP10c, KF12, KR06, KCO04, KL02, KFfKL11, LSJ06, Lac10, LK10, LLW06, LN01a, LLCL05, LMS96, LOR10, LS95, LV02, LL07, LYY07, LCY10, LHKŚ12, LY12, LMSS07, LGG97, LVP04, LM09, LLLL08, Li98, LZhL06, LZS^{+}08, LNW^{+}12, LY05, LAF^{+}10, Lin12, LN03b, LR05a, LCB97, LSL04, LL06, LBDW09, LW12, Loe74, LRH10, LOH05, LPK05, Lop06, LHK02, LZ05, Luk97, LMB08, LIH99, MSB12, MML97, MW91, Mal09, MS10a, MPB05, MSZ09, MVL11, MR91, MS00a, MZ04, MN02b, MW03, MB06, MS96b, MEB05, Men07, MSW02, MF11c, MT02b, Mey01, MMG^{+}01, MMS93, MV12a, MNS06, MS00b, MA05c, MNLM06, MMT08, MT99b, MF02, MS96c, ML95, MBJ^{+}07, MMH07, MAR05b, Mug97, MOTN07, MTNL12, MC98, MB87, MWJ97, MS11b, NWL06, NP09, NP12a, Nar01b, NZS^{+}06]. August [NSS^{+}07, NP12b, NKPH08, NOOS05, NT98, NR00, NF07, NW10, Nor04, NR08, NVSB^{+}07, Nos88, Nür03, NH03, OT08, Ohl93, OW09, Olt95, OTZ01, Ong05, OR08, OST11, Pal00a, PPCS12, PS09a, PKM^{+}10, Pau03, PPJRN^{+}10, PMT09, PMT10, PYLW08, PpWR96, Per11a, Per11b, PM96, PGBK04, PW03, Pfe06, PTM^{+}11, PG01, Pin95, PPBL12, PRJ95, PdA^{+}11, PT06a, PF07a, PRR94, PR97, PBW05, QSBC98, QBY^{+}12, Rab10, RWA06, Ral94, RM97a, RS99a, RS05, RBGS04, RWY97, RDPMA^{+}11a, RDPMA^{+}11b, RTB02, Rei95, RtTMP11, RMTW11, Rog11, RBSL06, Rov90, RV03, RSW12a, RG11, SN08, SNC12b, SW07, Sak91, Sak01, SGSV10, SGPP06, SS11b, Sal83, Sam76, STL11, Sch06, Sch09b, SW03a, SM06, Seh97, SPK01, SHX^{+}06, SW09b, SDH05, SS06b, Sho05, Sie86, SSAP05a, SSAP05b, SEL09]. August [SYP^{+}05, Smi12, SKY06, SLKL09, SF08c, SSU12, Sta82, ST12c, Sti94, ST01, SP12, STY^{+}07, Stu02, SBH^{+}04, SV12, TPB79, TMG10, TMdB11, TM99, TDS94, TFE^{+}09, THX^{+}07, TT05, TQYX11, TC98, TN04, TNY07, Tra04b, TCLH10, ÜÇA10, Vää99, VY01, VA07, VU05, Wag08, Wan01, WCO05a, WCO05b, WCO05c, Wan05b, WJ05, WFZL09, WZD11, WZJX12, Waś96, WR07, WS96, Wei97, WW02, Whi98, WH95, Wie99, WR09, Wil86, WTGA05, WSF08, WSJT09, WCJS10, WJ95, WMT96, YGGJ04, YEY04, YJ04, YJS^{+}06, YW06, YDW08, YSP^{+}10, YYS^{+}12, YI12, YLWY06, YKF^{+}06, Yi06, YAF^{+}02, YWG04a, YWG04b, YCB12, YY09, YZCW07, Yun02, ZEL04, ZL97, ZS00, ZGY04, ZSM^{+}09, ZZP^{+}10, ZCDT10, ZO10, ZJL06, ZNCF03, ZSY^{+}06, ZBH12, ZBG98, dCVK07, dGN12, dSK12, dSDL04, dSTP11, vBC08, vEGSW11, vLWH^{+}00]. August [vGH96]. August/September [PMT10]. Augustin [Her91, LK07]. Augustine [PM09, RVY05]. Augustl [LWM^{+}08]. AUSCRYPT [SP90, SZ93]. Aussois [GPM98, JRR03]. Austin [ACK00, BBA12, Gac02, HM04, HJ00, JS11a, KS09a, RL11b]. Australasian [BS02a, BSN06, BCD00, BG05, BN09b, DB98a, Li11a, MSS08, NL09, PH11a, PSNS99, PGD07, SNS03, SH10, SMS12, VPM97, VM01, WPV04, WR11, WZ08]. Australia [ADR07, Ami98, ATM09, AS98, AGT98, AJS00, ABD08, AOM07a, AOM07b, BB90, BS02a, BSN06, BTBP^{+}10a, BTBP^{+}10b, BKM08a, BHL11, BvLR04, BCD00, Boy01, BG05, BN09b, BS01b, Bra93, BkHJ^{+}04, CRW97, Che03b, CMS08a, CLLD09, CNA08, CKH10, CS12c, DSZ04, DDE04, DG96, DB98a, Dig04, DER94, Du00, DH08, EEM96, Ekl04, EL02b, FPRAW04, FL09, FK02, FG96, Foo99, GHM05, GF03, GKT10, GMO03, GHSW04, GN98, GMV12, GDG11, HENR03, HA02, HGZ05, HNF08, HNQ08, HSVZ10, IZ00, IPRO08, Joh97, KHJ05d, KjHGG11, KS02d, Kob06, KRH09, Kow01, KLC07, LOST04, LKZ^{+}08, Li11a, LABM^{+}12, LG05, Man09, MS03b, McC97, McK99, MS02a, MS03c, MB09, MS00b, MSS05a, MSS08, NL09, NRWI09, OT07, PWZ98, Pap95, PH11a, PHH10, PSN95, PS96b, PSNS99, POS00, PGD07]. Australia [Pie08, PKSS02, RAW07, SNS03, SNY06, Sat97, ShK06, SP90, SZ93, SA96, SYY^{+}08, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, Sta95, SH10, Stu08, SCB01, SLM07, SMS12, Tob80, VPM97, VM99, VM01, WPV04, WR11, WY04, WLP^{+}09, WMB10b, WMB10a, WKK98, WKD08, XCHZ11b, XCHZ11a, XLKZ12, Yao95, ZL96, ZL97, ZL98, ZS00, ZJ05, ZS07, ZPTI09, ZNCF03, ZSP^{+}04, ZYDL09, tHBP08, vSH06]. Australian [AS98, BB90, Foo99, GF03, KRH09, MS02a, OT07, PS96b, RAW07, Sat97, ShK06, SCB01, WY04, ZL96, ZL97, ZL98, ZJ05]. Australiia [WDKK07]. Austria [AI06, AdBB12, ALS^{+}05, Alp11, AHK07, AS07b, AKTT12, BBE01, BM03a, BBB^{+}07a, BP12a, BH06b, BS06b, BL05b, BCKR10, BKW09a, BBR^{+}10, BS02b, BKLR12, BMW10, Bos96, BS03b, Bou11, BNZ06, Buc85, BV94, BC04b, BW97, BFOY04, CCP96, CCF^{+}08, Cav85, CH10, CCG08, CKM^{+}09, Coi96, Con00, CBW99, Cro03, CHR10, CD12, DF10, DSB10, DKM^{+}05, DKN^{+}11, DR09, DB09, DKU05, DHO08, DPSU11, DFT09, DBH01, DFS06, EM03, ED06a, EBT11, EFT01, EK12, FM04, FHLP09, FHKQ12, FJL97, Fra00, FKBA12, FKM07, FK95a, FPRS06, GN90, GLM97, GN05, Gra05, GT93, GH93, GM05, HT09, HHMT12, HRdV11, HK94, HG00, HP06, HS97, Hic04, Hol07, Hol08, HM09, HS11, HRRY08, HNP12, HSL06, HS08c, HHG11, Hut94, JP07b]. Austria [JK05a, KR02a, KR04a, KJ09, KWC^{+}12, KKMW07, KS93, KLLP12, KBHSW06, KFMM06, KBH03, Kra03, KLN07, KKDV02, KSH05, KKH07, KS12d, KASS07, Kui92, KRS02, LHH10, LBP06a, LBP06b, LBP06c, LBP06d, LSTZ12a, LSTZ12b, LSK08, LS06b, MT09a, MPdSM07, MSZ09, Mau79, Mau91, MKZ02, MKZK06, MKZK08, MKZK10a, MKZK10b, MKPZ12a, MKPZ12b, MPPU11, Mio93, Mit05, Mos97, MB87, MJL02, MZ06b, PV07, PMT09, PSD12, PGG^{+}06, Pfi01, Pic86, PMD92, PMDA96, PMDK00, PPBL12, PDG92, QSBC98, RT07, RCT05, RK06, Rob06, SL12, SKMR94, SMH^{+}12, Sch10b, STT04, Ses06, Smo97, SS04, TR12, TQYX11, TBD12, TOL08, TdRM^{+}09, Val06, VHM07, VT97, Vol93, Vor08, Wii05b, Wim04, WSJT09, WL03, Win04, Win11, ZBW94, Zim92, ZVU99, vdV84]. Austrian [BBE01, Hol07, Hol08, HM09, HS11, KS93]. Authentication [KJR05, LLT^{+}04, MJ04, ZJ04a, BCB97, BS01a, KN03, LST^{+}05, TBJ02]. Authoring [Vaz99]. Automata [BCT02a, BMUV10, BI73, BJ01a, BMCCM11, CMZ99, DIMV09, DFMV10, DMV12, DS11c, ECB06, HŽ07, IY06, IR08, Man09, MVOF08, MR12c, RWY97, SB12a, SCH04c, UMN^{+}08, WLS89, WY98, CM03a, DIMV11, Dob81, DOSY05, FLS06, GP89, ID03, Pin89, Tau89, WW03, YP01, AS94, AGK^{+}10, ADG^{+}08a, ADG^{+}08b, AHS11a, AHS11b, AMSM^{+}09a, AMSM^{+}09b, ACJT07, AB78, ADCR89, BLPW03, BJ01a, Bra75, Bra85, BPSW06a, BPSW06b, CIM^{+}05, CMZ99, Cho88, CMPW12a, CMPW12b, DGMS97, Dia83, DKLS04, EK81b, FG95, GTW02, Kot86, Kui92, LSW98, LS88, LKC93, Loe74, Mau79, MM96, MAA91, NS82, NP85, OSvL01, Ott87, Par84, Pat90, Pup10, RWY97, SS77, WMR00, WTM^{+}02, WY98, vWN99, vLdB80, MB96]. Automated [Baa03, BK80, BSS11, Bun94, BD03, BR07c, CBC^{+}02, CYBF06, CY09, DV09, FS07a, Fri93, FKLY08, FTC^{+}10, Gan99, Kap92, KK98, LwCM00, LCY03, Lov82, LO88, McA00, McC97, MS96b, Nie05, PV00, Pfe07, Sch09b, Sho84, Sie86, Sti90, Vor02, Wan98, YEY04, YAF^{+}02, YTC^{+}07, YWRS11, YCB12, BR96b, BC98, BE01, BT06, Dal92, GHM05, GVM99, Lau93a, Pfe94, Vor92a, Vor93, Wil91, ABD08, AL11b, Baa03, BR04, Bec05, BSS11, BR07a, BC10c, BM11b, BH11, Bun94, CS00a, CCK^{+}08, CM12, Dyc00, EF02, FS06, Gal97b, Gan99, GWY99, GW09, GH10, GLN01, GZ06, GMS12, HW06b, Kap92, KK98, LR09, MP03b, McA00, MP96, McC97, MS96b, Mur99, NYHO07, Nay95, Nie05, Oli07, PT05, Pfe07, RGW01, Sch09b, SNRG11, SZ11, Vor02, Wan98, Wan04, Win04]. Automated [de 98]. Automatentheorie [BI73]. Automatic [Kai11, Urs02, Bre97, Gor02, Bur97, Fra96, LZ02, Ral81, Sif90]. Automation [Sch98b]. Automotive [BKM06, BKM08b, BKM06]. Autonomic [BFvdM10, GTP12, HRR^{+}06, SGD09, vdMBD08, CYM^{+}11, GPAS^{+}06, HJMP06, NRWI09, PMÓ06, RJS^{+}08, Smi05, SS06d, XYM^{+}07, XBS^{+}10, YJMU06]. Autonomous [ABH^{+}10, BB07, Bee02, Bv97, CNP^{+}08, CW94a, CM95, CCBW11, ENO12, EEM^{+}11, GANACBR10, GANACL11, GANACBdCdV12, GB99, GAM^{+}11, Har11b, HS08b, HSP^{+}12, NABK12, PM96, SP09, SNČ^{+}12a, Sch10a, SD10, VP96, Oss99, TP97, von92, GLS05, GZSC07, Han02, KKGK11]. Autonomy [Kim05b, NRW04]. Autrans [FD03]. Availability [MRvM07, NMPM08, PRT06, QBY^{+}12, MRK05, MNS05, TQYX11]. AVBPA [BCB97, KJR05, BS01a, KN03]. Aveiro [CK12b, CK12c, LLMR09, OMMSP06, TdLdOQ08]. Avi [Gol11b]. Avigad [Gol11b]. Avignon [ABG96, BST01, BCC^{+}12, CT07]. Aviv [TCYZ12]. AVYTAT [MDH10c]. Awaji [TNI09]. Aware [EEM^{+}11, JvdA09, MKPV05, KKM04, Leh11, FV01, FV03, FV04, FV05]. Awareness [CQSS09, HKS06, HSS07, RTB02, SLP05]. away [VP96]. AWCC [CL04]. AWeSOMe [MTH05, MTH06a, MTH06b, MTH07a, MTH07b, MTH08]. AWIC [SKN05, FMC04, MSS03a]. AWOC [Rei88]. AWSN [THX^{+}07]. AxGrids [Dik04]. Axiomatising [Gol82]. Axioms [Knu92]. Azores [FLMG10, MZF94].

B [BBBB08, DFG^{+}12, FGK^{+}10, JK06, TKHS05, BG93b, Ber98, Ber02b, BBKW03, BDGK00, JK06]. B.C [JS94]. B’98 [Ber98]. Back [GCSR07, Wil01b]. Bad [BMS04a, BH77, BRR87a, BRR87b, CER79, Dit88, DPP97, Gab97, HS76, HNW04, Jan88, Klu97, KK95, LK99, Len93, MT86, Nol81, dRLP98]. Bahia [MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, KMS10]. Bains [KR02b]. Bakuriani [BLSS11]. balancing [BGW93, HWP01, MNW08]. Balatonfüred [DKP00, LRS03, Rue93]. Bali [BMC08]. Baltic [BB91]. Baltimore [FNY93, GM99, Lam96, Par02, SBF08]. Bamberg [Kat99]. Bandung [ABK05]. Banff [A^{+}08, BKO04, CYM^{+}11, CKK^{+}09, DGST09, HYMZ11, HNPP07, JPBT05, Lin07, Lin12, LOH05, MH04, STS04, TBKO10, YRWS11, ZY01, vdPS08]. Bangalore [AYV^{+}11, BP04b, BL07b, BB00b, GM10, HMV01, JM07, JS84, Mad89, MMV06, Nor90, Ram08, RPBP06, RD12b, SPBP08, SKS02, Thi95, VKV00]. Bangkok [AAW04, CYCL09, CJ05, DA08, FNPW05, GGS09, KKMN07, LLC09a, LLC09b, MWK^{+}09, SMEGA10, TKCH09, TNA^{+}10]. Bangladesh [iNR08, RF10, RiN12]. Banská [Rov90]. Banz [GS88]. Banz/Staffelstein [GS88]. BAOSW [NSM07]. Barbados [CSN14, DG09b]. Barbara [AH02, Bel00, Bon03, Bri93, Buc90, Cop95, Des94, Dwo06, FR95a, Fer96, Fra04, GF12b, Hal09, HT06a, ID03, ID06, Kal97, Kil01, Kob96, Kra98, MS10a, Men07, Pal00b, Rab10, RMDS10, Rog11, SNC12b, Sho05, SSM08, Sti94, Wag08, Wie99, Yun02]. Barcelona [AHSC11, ADL09, BYdVZ^{+}12, BLS06, BBGS10a, BBGS10b, BBGS10c, BHJKPG06, BMSMS04, BGP11, BBRS05, CPS05, CA08, Cho12, CNmWH^{+}05, CEKR02, Cra08, CRW^{+}12, DH05, Dia83, DO89a, DO89b, DS96, DJRZ06, DGDM^{+}11, DFT04, DLM99, Due04, FZJE^{+}05, FZJE^{+}06, FMO05, FRSH05, GG05a, God05b, GM04b, HT05, JP04a, KN09a, KMD02, LO11, LTÅ^{+}11, LV06, LSYO08, LK04, LP00, LRS98, MMD05, MOT12, NDB94, Nao07, OP97, OS09, PLMV12, Sch04b, SS05b, SS06a, SSJR10, Ser12, SPSV^{+}04, SPC^{+}08, SQL10, Sur02, TLJ06, TN04, VDPD96, VBS^{+}10, Vit95, WR01, Wal04b, WMS04, WKS^{+}10, Zui99]. Bari [AE05, Esp01, ERMS06, vLS91]. Bariloche [HOC07]. Barriers [BBLT06]. Base [CW07, TDG91, Bla80, DS01a, HS76, Mat91, Yao82, YK82]. Based [ALS^{+}11, AW01, ABB99, ABMH08, ADM^{+}08, AB03, BdB08, Ber02a, BJSW09, BM10, BJ07, BMYZ91, BMYZ95, BRM06, BSS00, Bru04, BL04b, CMSM^{+}10, CW11, CUDL02, CR04a, CMS11, CÓ11, CS12c, CP02, DYCD09, DG02, DY10, ERRW03, ELLP12, FCLL02, FdS12, GW06a, HO03, Hil06, HRT^{+}03, HRR^{+}06, HKH10, JS02, KJR05, KP04, LLLL08, LPC^{+}11, LZS^{+}08, LMP04, LLL06, LSW^{+}10, LCY^{+}11, MPP08, MW09, MdPA98, MAR05b, MGSM12, NNS03, Nar01a, NHJ04c, NKPH08, OPS11, PPPM10, Pet02, PLK^{+}12, Qui04, RW11, RHD08, RBGG06, SPG10, SVS08, SLKL09, THR03, Ull08, VY93, WAS06, WR07, WAW01, WL04, ZZ04b, ZNCF03, Aus93, BHGP02, Ber99, BCB97, BS01a, BP01, Bos01, BJ01d, CRW97, CKK^{+}09, CBN99, CE00, DLT05, DN04, DC01, Eta05, FG04b, GH92, Hag02, Hal96]. based [HV03, HKM95, HRTR05, JL95, KN03, Lau01, LLCL05, Leu03, LMA09, LSL04, MMD05, MSY00, Pad99, PWZ98, PS97a, Pup99, RCA90, Ras01, SCG98, SF96, SC98, SZ05, Vei03, Wag94, Wat95, WP90, WAR94b, YBM87, dBBC02, AW12, Alp11, AT07, APN08, AHJ07a, AHJ07b, AHJ07c, BGP11, Bee02, BM10, BHVV09, BGJ11, BRM06, BJK^{+}05, CV04a, CPV03, CSR08, Col07, CR01b, CSSW04, DS09a, De 10b, DP10a, DGWB09, EBT07, EV07, GHJ06a, GHJ06b, GHJ06c, GP08a, GKL^{+}11, GHC^{+}06, GRP10, HENR03, Han09, IS12c, JFT11, JMO10, KHJ05a, KHJ05b, KHJ05c, Kin08, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, LBSBW98, LPH09, LHJ08a, LHJ08b, LHJ08c, MW09, MAR05b, Nar01a, NHJ04a, NHJ04b, PHJ03b, PHJ03a, Pau01, PH11b, Pue07, RW11, SCHS07, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, SW09b, SA06]. Based [TOOO07, TBE08, VRHJ09a, VRHJ09b, Vid12, VSMS12, WR07, ZZ04b, AW01, ABB99, AB03, ABGP05, BR97, Bos00b, Bra99, BGY96, BV05b, CNY95, DM01b, EW03, FL00, Fle99, FP05, GNR94, HCS^{+}05, KHJ05d, LP97, PS98a, SBD03, SL99a, TWN92, TEB09, VA95]. Bases [ST08, ST11, BMYZ91, Hil78]. Basic [RR98b, SKS^{+}09, Sch97a]. basics [Pod95]. basis [FC93, FK95a, HF97, LCS91]. Bastad [DNS95, JM95]. Bath [Fit93, HL04, PCG^{+}05b, SSAP05a, SSAP05b, Wal84]. Baton [ANS^{+}09a, ANS^{+}09b]. Batumi [tCZ07]. Bay [BW05b, Bru06b, DDS12, Mon01, SV05]. BC [AW01, BS11a, HM05a, IS11b, KKGK11, KC11a, KC11b, PSS^{+}10a, TDK^{+}10, BM02b, HV98, MN98]. BCS [CGvR02]. BCS-IRSG [CGvR02]. BDA [SB12b]. Beach [AN09, Baa03, BA07, CC07, DG09b, HC05, HM96, Jen01, KP04, KS07b, LN09]. Beaune [IHQ96]. Beauport [JP11a]. Beekbergen [dBdRR93]. Behavior [AHMT08, BSG03, BSPB07, Dad01, DGG^{+}10, NBC^{+}06, PBSB09, SGSV10, SL11, SdSMO12, ZBH12, BSPB07, Tu99]. Behavioral [CSM10, SYNC11, YGE12]. Behavioral-Cultural [SYNC11]. Behaviors [MSY08]. Behaviours [EFZKM07]. behind [CW02]. Beijing [ACL12, Ayk07a, Ayk07b, BW03a, BYLJ11, CCL06, CIW06, CPR^{+}10, CPSA10, CZG^{+}08, CWZ^{+}11a, CY10, CY11a, CYZ12, CvSW04, Dai07, DZ94, DFW^{+}11, Duf07, FLY06, GWY99, GML^{+}08a, GHSY06, GZ06, HOQ97, HTW02, Har07, HLKX12, HPR05, HM12c, HYP^{+}09, IS12c, JZR09, Jac07c, Jac07d, Jac07b, Jac07a, JNPV10a, JNPV10b, JNPV10c, JRJ05, LST^{+}05, LBO05, LEP^{+}10, LYL06, LSL04, LWYN09, MCA08, MDY^{+}10, Mei08, MSW02, MLTC11, MSG06, MZWZ05, Nob12, Oho03, OiDP98, OW07, PKM^{+}10, PFCS05, QMLW05, QMW09, QSWL11, SR07, Sch07, SLH05, SXC04, SvADS07a, SvADS07b, SvADS07c, SvADS07d, SLC01, Shu07, SS07c, SS07b, SLSW12, Ste07d, Ste07b, Ste07c, SZT^{+}08a, SZT^{+}08b, SLCT11, TSG00, TST10a, TST10b, TKCW11a, TKCW11b, TTP05, TCLH10, VHM07, WYSS10, WS12, XCR11, YJ04, YX04, YC11, YGL^{+}10, YLL09, ZLJ05, ZOCJ07, ZGT05, ZZ99, ZLY^{+}07a, ZLLC09, ZOM05, ZY10]. Beijing [ZF05, vdHXMW08]. Being [IdKM06]. Beja [PA03b]. Belfast [BH06c, BW10b, BW01, BLS02, EGMW11, KKM^{+}11, Liu11, MA03a, WR07]. Belgium [AD11, BDE12, BPV04, Bjo87, BTPPS05, BTPPS06, BTKP^{+}11, BG08, BW92, BD03, CW11, CDY^{+}11, CLS07a, CS01b, DGM08a, DGM08b, DR02a, DKS^{+}07, DLNU11, De 10a, DMB^{+}11, DW12, DB04, DB98b, DY10, DIP12, DDS02, DBB^{+}04, DGB^{+}06, DBB^{+}08b, DBD^{+}10b, DBB^{+}12b, Dun09, EPVW10, FS07c, FSP09, Fuc98, GANACBdCdV12, GPV88, IS08, JP05a, JDL11, Jul98, KR10b, Le 94, LCHS96, Mag98, MBH^{+}09, MRZ09, PVW10, PVV05, Par84, PNPP12, PGV93, Pre95, PR98a, Pre00, Pro11, PS12c, QV90, Qui98, QS00, SSv94, SBH07, SBH09, Tri98, TWSN07, UPB07, VW07, Van05, Wol95]. Belgrade [LPR12]. Belief [KI01]. Bellare [Gol11b]. Benchmark [KS09a]. Benchmarking [KS09a, NP09]. Benchmarks [KGS08, BGS^{+}03]. BenchmarX [CLLD09, YMY^{+}10]. Bend [CDCW09]. benefit [GW96, NDB94]. Benicassim [MdPA98, dPMA98]. Berder [GCK06]. Berg [BHK01, GW03, VvS99, vEGSW11]. Bergen [ÅBE12, AR04, DFD04, Fom06, GKS^{+}07, Hal00, JK04, Kap10, MMH07, PSO05, RS00, SMMG01, Ytr06]. Berger [GK90]. Berkeley [CJRT05, DJNR09, Fed01, Fer98, GJRT01, HS98a, HJRS99, JSW03, KS03a, MS12a, Nor97, PDML11, PB92, Sel86]. Berlin [AKK04, AB11, ASS10, ABB10b, AH03, AH10, BE11, Bel03, BS03a, BHMS07, BGPTF10, BLB^{+}03, Bis02, BH92, Bla90, BE07, BFH98, BTT12, BL06d, BMP96, CZ02, Ceb08, Cov99, CRS00, DH07a, DLN10, DW07b, DA01, DS07, EV96, EHP94, Ehr85a, Ehr85b, Ern07, FMNS06, FSS06a, FSS06b, GL97, GBMS11, GS00b, Ham89, HWP01, HPP06, HW07, HS98d, JK05b, KWW04, KO11, KHW03, LM08a, Mai00, MNS05, Moh91, Moh97, Nar01b, OZ01, OTZ01, PW92a, PW92b, PBH03, Per10a, Per12a, Per12b, REF^{+}06, RPZ97, SS83, SH08, SV97, SPZ90, Sie75, Smi05, Smo00, Som96, Ste01, TY11, Tiu00, Wat00, WLS89, Wol07, ZSP10].

Bermuda [Bla03b]. Bern [BN91a, BCF06, BM11b]. Bernardo [dRCVT11]. Bernd [Hei12]. Berne [Bor92]. Bernhard [DKS12]. Bernried [BR98, TS87]. Bertinora [BC04a]. Bertinoro [ABBC10, ABDW10, AC11, ABD03, AM05, BI03, BB05b, BH07, BDZ08, BPZ09, BI11, BCP12, BL03, CB06, DFUR11, DPZ12, DO09, DO11, FFL^{+}12, KM08a, Lag05, LPR08, Sch03a]. Besancon [JK06]. Beth [CGMQ08]. Bethesda [CSM10, RBF06]. Bethlehem [BL05a]. Betrieb [Lei74]. Betriebsprogrammierung [HS74]. Between [FSdlPT11a, FSdlPT11b, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, Gol11b, Mit06, MHS10, Pad99]. Beyond [BDFM12, BL10b, DKN12, KS02a, KS02b, MSWZ06, MHD09, PS11a, VSTN06, BM99b, FR99b, KN91, Kun84, Sch93a, She03, SLPM96, SHM^{+}10]. Bhubaneshwar [NO11]. Bhubaneswar [Cha05a, GM04a, JM10, MCK^{+}06, PDSN12, RR12, YAKM^{+}12]. BI [CVLL12, DMB^{+}11, HLCZ11, NSM07, YSP^{+}10, ZTTY12, ZLLC09, BGHS05, GHSW04, KBHSW06, KHSM^{+}08]. Bi-Conference [BGHS05, KBHSW06, KHSM^{+}08, GHSW04]. Bialowieza [ABT04]. biannual [Pup99]. Biarritz [CGP02]. bibliography [BKL^{+}91, DJL88]. BICoB [Raj09]. BICS [ZHLW12]. BIDM [IHX^{+}08]. Bielefeld [EJM10, KW10, RBW95, WF98, WK08]. Biennial [EJM10, Ham00, KKKS11, McC96, MN98, SM01, ADB^{+}05]. Big [DKC10, WZ03, SB12b]. Bilbao [AFGvE10, BHQ10a, BHQ10b, FP05, HMT10a, KLS10, KLP10, PMT10, RW06, vdL02]. bilinear [De 87]. Binding [DHM99, JM78]. Bio [BKLP11, BKLR12, KLP10, LYCV08, MSA99b, MP01b, CSPC09, HGPH12, LFIM07, LYCV08, MÁ07a]. Bio- [BKLP11, BKLR12, KLP10]. Bio-Inspired [LYCV08, MSA99b, MP01b, CSPC09, HGPH12, LFIM07, LYCV08, MÁ07a]. BIOADIT [IMK06]. BioAW [MJ04]. biocomputation [BE95]. BioDM [LYT06]. Biohybrid [PLMV12]. BIOID [SJDT08, VDD^{+}11]. BioID MultiComm [FOGE^{+}09]. Bioinformatics [BCM08, BP03, BVAT12, BM06, CNA08, CL08b, DKY06, DTMH10, FMS10, GVB12, GH07a, GG02b, GPP09, Hof97, HLI06a, JK04, KSG^{+}09, LMSS07, LWRdR11, MM08, MTV09, MPT10, MS10e, ORB^{+}09, PRG09, PRG10, PRG11, PS11b, RWA06, RSV07, RT12, RL11a, SW07, SW09a, SVA05, SKSA12, TWSN07, dSK12, dSTP11, GM01, MMR07, MC05, BMSW12, BGPR10, CWZ11b, FN12a, HW07, Hof97, MZ07b, MSZ08, MNZ09, Raj09]. Bioinspired [CPS05, FSdlPT11b, MÁ05a, MFÁ^{+}09a]. BioLink [BS10]. Biological [ADM11, DKOZ05a, BMSMS04, MCKB06, MMMD97, OMMSP06]. Biologically [BLPW02, MRC12, IMW04, IMK06, LBP00]. Biology [AHK07, AGI^{+}06, BS11a, Bat09, BCM08, Ber10, BDZ08, BS10, CG07, Cho12, DG09a, DSS^{+}07, EIRW06, FMS10, Fis08, GPP09, HU08, HRRY08, HNP12, HSH08, IB07, KLZ07b, KRH09, LMSS07, MMK^{+}05, MMC05, PMGO05, Pri05, PZ05, Pri06, PCE06, PHPL06, PP06, PIMN06, Pri07, Pri08, PDAN08, PBP09, PBG^{+}10, PBPdV11, PPdV12, Raj09, SW07, SVA05, SH07, TTM10, VW08, dSK12, dSTP11, CER79, DS05a, GS01, HML97, Liu01, Mil00, Pri03a, SMPU98, THT03]. Biomechanics [Tu99]. Biomedical [BHVV09, HKW^{+}11b, LYT06, SKK04, HK96, ZLJ05, BE08, BC10a, DCFR12, FDL10, GMV03, HS06, PLG06]. Biomedicine [PK12]. Biometric [KJR05, LLT^{+}04, BCB97, BS01a, KN03, LST^{+}05, FOGE^{+}09, MJ04, SLCT11, TBJ02, ZJ04a]. Biometrics [KZ10a, LL07, TN09, Zha07, ZS10, TBG05, ZJ06, SJDT08, VDD^{+}11]. Biomimetic [PLMV12, WPE05]. BIOS [CPGL05]. BioSecure [ZCRL08]. Biosecurity [ZCRL08]. Biostatistics [BVAT12, MTV09, MPT10, RL11a]. Biosurveillance [ZGK^{+}07, ZCRL08]. Biotechnology [PS07a, PS08a, Per05]. BIOWIRE [LYCV08]. BIRD [HW07]. Birmingham [ACR^{+}08, CRW08, FKBP11, Sex09, YBL^{+}04, YTC^{+}07]. BIRTE [BCDN07]. Birthday [DKN12, Mah04, MRU12, AG09, ADM11, AAN09, ACÇ^{+}12, ADR08, BS11b, BDR10, BDFM12, BKT12, CRS11, CLKK07, CS12b, DDM08, DKS12, EMO10, ELS^{+}10, FJM06, GKKW09, HL05, Hei12, JL11, JPR04, KK11c, Len04, LLM09, MM09a, Nac12, Pal09, PQR11, RSB07, SPG10, SW10, ACG^{+}07, Der04, HNR05, HS05b, KMPR04, KMO^{+}05, MvOvRdV05]. Birthdays [JLW07a]. BIS [Abr07]. Bisociative [Ber12]. BITCOIN [CSN14]. Bjørner [JLW07a]. Blacksburg [CT99]. Blanca [KMS10]. Bled [BGMST09a, BGMST09b, DF99, PLC11, RtTM12]. Blended [LWM^{+}08]. Blind [PP04, REPH06]. BlogTalk [BBK^{+}11]. Blois [BMCCM11]. BMCV [BLPW02, LBP00]. BNCOD [BH06c, CK07, FGB11, GJS08, JNS05a, Mac12, Sex09, Bow94, ENP02, Emb98, GK95, GL92, GW93, JLY03, LJ00, MK96, Rea01, Sma97, WM04]. Boaz [ADR08]. Boca [BTMI11, BHLP07]. Bochum [AV07, LSW08, ÖRH12, RBG09, von96]. Bohinj [SGG06]. boiler [ABL96]. Bolc [MM09a]. Bologna [BM96, BZ09, CNY95, CW06, CR04b, DGMS97, DG09a, FBKS03, GOW03, GW06a, LM93, LM00, LS10a, LMP04, LMT10, MK03b, PSP^{+}02, Rol97, SBD03, TP94, Ver08]. Boltenhagen [BL01]. Bolzano [BGPW05]. Bombay [Bha95, RCA90, Shy93]. Bonaire [BDC12, Ker12]. Bonas [BCM00]. Bonn [AWA^{+}11a, AWA^{+}11b, BH06a, BI73, BHLvdP07, BCC99, CBBS09, DMWW88, DRS05, GO94, GO96, GMV07, KP05, KJ10, LK09b, MW11, MTWD06, Ohl93, RB10]. book [BB96]. Boolean [Tha84, Tha81]. Boppard [HMSU04, HU05, HSTU99]. Bordeaux [AGK^{+}10, ADB^{+}12a, ADB^{+}12b, BSBM05, BTPPS09, BLV02, CW88, FLP12a, JNR11a, JNR11b, Kla12, Les87, NvRSS11, Sch10a, ST11, Shv06, dGEK11]. Börger [AG09]. Borgholm [Kar83]. Boris [ADR08]. Borovets [BDGN07, Dim03, DDK11, PCA02]. Borrowdale [Tho11]. Borrowing [KRH09]. Boston [AP04b, ACHH12, CL01, Con97, DM01b, Dov05, GJRR08, KK00, LBDW09, NLS08, Nor04, PZK07, PC00]. Botik [MT89]. Boulder [EM02, HS03]. boundaries [Vel94b]. boundary [Chi79]. boundary-value [Chi79]. Bounded [Ram96]. Bournemouth [GHJ06a, GHJ06b, GHJ06c]. box [CKPB07]. Bozapalidis [KR11]. BP [ALS^{+}05, RBsC^{+}06]. BP-UML [ALS^{+}05, RBsC^{+}06]. BPD [BH06d, ED06a, tHBP08]. BPI [BH06d, ED06a, tHBP08]. BPM [ADR07, BGK12, BH06d, DEKR09, DRS08, DFS06, ED06a, HMT10c, RMTW11, tHBP08, vdABCC05, DPW04, vdAHW03]. BPRM [BH06d]. BPS [BH06d]. BR [CDFP11, Nic86]. Braga [De 07b, DL07, FLVS11, GH07b, KO07b, KMR12, LSV06, LVS08, LP10a, LV07b, Sei07, SA12, SHO99]. Brain [LSIT11, MSY08, MRSV07, SKS^{+}09, YLS^{+}12, ZHLW12, ZLY^{+}07a, DDFM05, HLCZ11, YSP^{+}10, ZTTY12, ZLLC09]. Brain-Like [SKS^{+}09]. Brains [BSPB07, Roc92]. Brakerski [Gol11b]. Branching [Pup10, Mei89, dBdRR89]. Brandenburg [VDD^{+}11]. Brasilia [Cur09, DdQ10, Esc10, dSTP11, Tre09]. Bratislava [GRW86, Jay99, KM11, MN02a, NR00, PR97, RV03, RSW12a]. Braunschweig [Lei74]. Brazil [AJRB09, Agu12, BPAB07a, BPAB07b, B^{+}12a, BL04a, BCM08, BR02, BC10b, BCSW10, BK96, BC09b, CSW06, CDGW10, CG02, CEM^{+}06, dRCVT11, Cur09, DSS11, DdQ10, DPC^{+}07, DLdS04, DMB^{+}07, ERK07, ES03, Esc10, FMS10, FLS05, GN12b, GPP09, Hae99, HP09, dCJB12, KKM04, LBNF08, LCD^{+}09, LdQ07, LVP04, LSHB07, LLO08, LM98c, MSAdAS05, MEB05, MS00c, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, MB97, NdMO03, OW09, PBK^{+}10, PSM10, RM04a, SW07, SVA05, SCR06, SM11b, Sim92, SMK01, TDWG11, Tre09, V^{+}06, VGB11, WC95, YCB12, ZdC08, ZZTG07, dCVK07, dLWC03, dO98, dSK12, dSDL04, dSTP11]. Brazilian [B^{+}12a, BL04a, BCM08, BR02, BK96, dRCVT11, DSS11, FMS10, GN12b, GPP09, dCJB12, MS00c, MB97, OW09, SW07, SVA05, SCR06, SM11b, WC95, ZdC08, dO98, dSK12, dSTP11]. breaking [VP96]. Breast [MBG12]. Bremen [BKLM07, DVM^{+}11, EKR91, EEKR12, FKS07, HS08b, HMM12, HG98, JCC^{+}12, Koh06a, MOH^{+}12, Pfe07]. Brescello [FdCLN^{+}11]. Brescia [Squ09]. Bressanone [HL10, TFE^{+}09]. Bressanone/Brixen [HL10]. Brest [All96, CLB11, DEM96, KK09b]. Brian [JL11]. Bridge [Mit06]. Bridging [MHS10, PST11, ZRP^{+}09]. Brighton [AM91b, AM91a, BD98, Dav91, Fog96, Gol94]. Bringing [CCK95]. Brisbane [ADR07, AS98, AOM07a, AOM07b, BCD00, BG05, BN09b, BkHJ^{+}04, CLLD09, DG96, DB98a, DH08, GHM05, GKT10, LABM^{+}12, NRWI09, WKD08, ZL98, ZPTI09, ZSP^{+}04, ZYDL09, tHBP08]. Bristol [BM03c, DFR02, FDC12a, FDC12b, HSP^{+}12, SLL01, TG00]. British [BH06c, Bow94, CK07, CMQ96, DR06, ENP02, Emb98, FGB11, Fra01, Fra99, GK95, GL92, GJS08, GW93, Hir97, Hir98, JNS05a, JLY03, LJ00, Mac12, MK96, Rea01, Sex09, Sma97, Syv02, WM04]. Brixen [HL10, TFE^{+}09]. Brixen-Bressanone [TFE^{+}09]. Brno [BTPP^{+}12, BJKK02, BR07c, CCMR10, DV10, GLM93, GGS99, HK10b, HKM^{+}11, KP10c, PBRB08, SKP00, SKP02, SKP04, SKP06, SHKP08, SHKP10, SHKP12, ZBG98]. Broadband [CCK95, KG04, KMN94, Kim05a, PPF^{+}00, RMV96, Pla96, RMV96]. Broadening [CB12, DS01a]. broader [She03]. Brockenhurst [BEP04]. Brook [Cle92, Cow77]. Brooklyn [Par85]. Bruges [GPV88, Pre00]. Brunswick [AH96, AKS09, GKL09, Gan96, MS96b, ZGK^{+}07]. Brussels [Bjo87, DMB^{+}11, DB98b, DDS02, DBB^{+}04, DGB^{+}06, DBB^{+}08b, DBD^{+}10b, DBB^{+}12b, EPVW10, JDL11, Jul98, LCHS96, Mag98, MBH^{+}09, PVW10, SBH07, SBH09, VW07]. BSB [BCM08, FMS10, GPP09, SW07, SVA05, dSK12, dSTP11]. BSDIA’97 [MB97]. BT [VMvN12]. Buchmann [SPG10]. Budapest [Ama08a, AFR^{+}04, BBvRL10, BDG04, BDG^{+}12b, BP03, BBT10, CGL^{+}09, CS10, CM09b, CTE^{+}06, CK93, CVÉ07, DKP05, DZ03, DP10a, DH06, Dro08, EVV^{+}11, FI08, FGTZ08, Fro02, GGS11, GR06a, GW12, Hen08, Hor06, HT06b, HZB07, HPZ10, JSO04, JFBH08, KKS11a, KKS11b, Knu01, KFM07, KVMH09, KKD04, Leu08, MML97, MRPM10, MVA01, MWJ97, OV09, PAV^{+}10, PT08a, PJM^{+}08, PVP05, RR08, RAM05, RBC^{+}06, SVV12, SV12, TT10, Vog01, Wei97, ZHP12]. Budejovice [MM03]. Budva [ABMS06]. Buenos [AMGJ12, CN05, FC04, OdQ12]. Buffalo [BBH^{+}10, BBH08]. Bug [Lib07]. Building [BSW12, Nag96]. buildings [SKB98, Str99]. Bulgaria [BDGN07, CKD11, CD00a, DD10, Dim03, DDK11, Dji89, DPT08, ED06b, Giu98, INR07, LVW05, LP10b, LMWY04, LMW06, LMW08, LMW10, LMW12, LNSS11, MWY01, MVW09, OF15, PCA02, RA12, Sco02, VWY97, VWY01]. Bundy [Koh06b]. Burg [Bet83, ST00]. Burgos [CYBF06, CAP08, CY09]. Burlington [Bue04]. Burnaby [KKGK11, KC11a, KC11b]. Busan [ACN^{+}07, HYPY10a, HYPY10b, KG04, KT06, KKH09, gLPZ^{+}12a, gLPZ^{+}12b, RL06, YY09, YYH^{+}12]. Business [ANF08, ABR09, BKZ05, BHM^{+}02, BFOY04, FKB02, FHFL06, FHLP09, FHKQ12, FKL08, FLP11, HTMC09, KLP04, KLS10, LPT07, Lei74, LPK05, Rei03, TQYX11, VA07, LMT00, Lou94, SCKR03, Abr07, ADR07, ADM^{+}12, BGK12, BH06d, BCDN07, DEKR09, DPW04, DRS08, DFS06, ED06a, HMT10c, RMTW11, tHBP08, vdADO00, vdAHW03, vdABCC05]. Búzios [LBNF08]. BVAI [DDFM05, MRSV07]. Bystrica [Rov90].

C.I.M.E [GI03]. CA [AD05, Abe06, ADJ^{+}01a, ADJ^{+}01b, Ayk09, BPCS06, BA07, BHHJ11, BC07, BFMS08, BKM06, BKM08b, BG07b, CS11, CS05a, CC07, CJRT05, CGS07, DG07b, DJNR09, DLW06, DF06, Duf09, DG05b, Dun12, EJM10, EW05, EIRW06, FV04, Fer96, Fis09, GAA06, God05a, GZ09, GGHS07, GYA^{+}06, GF12b, Hal09, Har09, HMP08, HC05, HT06a, HW09, HW08b, ID06, IB07, Jac09c, Jac09d, Jac09a, Jac09b, JT09, Kar09, KS12b, Kia11, KS07b, KSHM09, KP15, KS10b, Kur09, LSgL^{+}06, LN09, Lin14, LR05b, MS12a, MWYF10, MT09c, Mal08, MRY^{+}11, MZC06, Men05, Men07, NNPW12, OPKR05, OZ09, PM08, PDML11, PS09a, Pie10, Poi06, PISW05, Rab10, RMDS10, Rei09, Rog11, Sab10, SNC12b, ŞS05a, SS09a, STL11, SEG09, SW09b, Shu09, SS09c, SSM08, SBM11, SH07, Ste09b]. CA [Ste09d, Ste09c, TD07, VH04, VS05, Wag08, XL11, ZSM^{+}09, dKIM^{+}07, AH02, Bla03a, CF00a, Cha02b, Cou03, CG94, DS01a, DS03, Ell95, ET01b, Fed01, FR95a, GHR89, GJRT01, GK02, GW98, HPV00, HK01, HJRS99, HH03, ID03, IL01, Jen01, Joy03, KS03a, KKP02, KP04, KMSS02, LSW02, LMW01, MS00a, MTO99, Mon01, MS02c, Mue01, Nac01, Oka04, Pal00b, Pre02, PB92, RLL04, Ric02, SW04, SC96, SWB03, TG02, ZZ04a, ZG03, vEDA04]. CAAN [Erl06, JP07a, LOH05]. CAAP [AM91b, Arn90, BD97b, DN88, DO89a, Ehr85b, EEK^{+}87, FZ86, GJ93, Kir96, MNS95, Rao92]. CAAP/FASE [BD97b, GJ93, MNS95]. CAAP’94 [Tis94]. Cabos [CL10]. Cachan [CG88, FJ92]. Cachan-Paris [CG88]. Caching [CvSW04]. CAD [Enc80]. CAD-systems [Enc80]. CADE [BSS11, McC97, Nie05, Pfe07, Sch09b, MS96b, Baa03, Bun94, Gan99, Kap92, KK98, McA00, Vor02]. CADE-11 [Kap92]. CADE-12 [Bun94]. Cade-13 [MS96b]. CADE-14 [McC97]. CADE-15 [KK98]. CADE-17 [McA00]. CADE-18 [Vor02]. CADE-19 [Baa03]. CADE-20 [Nie05]. CADE-21 [Pfe07]. CADE-22 [Sch09b]. CADE-23 [BSS11]. Cadenabbia [NDFMS97]. Caen [EMW94]. CAEPIA [BCC07, LGM11, MOBS06, MMG10, CUdlC04]. Caesarea [HG02]. Cagliari [KR00b, RK02, RKW04, RV05, TAS05]. CAI [BR07b, BR09, LMCH12, Win11]. Caianiello [GG98]. CAIP [GP05, JP09a, KKH07, RDPMA^{+}11a, RDPMA^{+}11b, CK93, HS95c, PW03, Ska01]. CAIP’97 [SDP97]. CAIP’99 [SL99b]. Cairns [AGT98, CRW97, FG96, HA02, Sta95, WY04, ZL97]. Cairo [EKR10, Ima99, SE10]. CAISE [BDGM05, PF05, BL08, DP06, KOS07, MR11, Per10b, RFBW12, vEGW09, ABS91, Bre03, DGN01, EM03, ILR95, JO99, Lou92, PMWO02, RBC93, SSB90, WB00, WBW94, BHM^{+}02, CMV96, PS04b]. CAiSE’97 [OP97]. CAiSE’98 [PT98]. Cala [AS06]. CaLC [Sil01]. CALCO [CKC11, FHRR05, KLT09, MMH07]. Calculemus [ACR^{+}08, ACD^{+}10, CBC^{+}02, CDCW09, DFUR11, JCC^{+}12, KKMW07]. Calculi [Cur09, Ong11, Roc07, Abr01, BG93b, DH97, DCP95, Gir99, Hof03b, Urz05, Bau98].

Calculus [GLW07, Boh75, LT01, Mil80, Tha81]. Calcutta [BKS99, DB02, PS02a, PKM04, RO00b]. Caldes [EO94]. Calgary [BFSN10, CS02a, JRSN09, SN08, ZEL04]. California [Ban92, Bel00, Bon03, Bri93, Buc90, Cop95, Des94, Dil94, Dwo06, EG97, Fer98, Fra04, HS98a, HM96, IR08, Ish97, JSW03, Kal97, Kil01, Kob96, Kra98, LG94, NW97, Nor97, Por98, Ran93, Seh97, Sel86, Sho84, Sho05, Sle92, Sti94, Wie99, Wil99, Yun02]. CALISCE [DF96]. Callaway [HNM02]. Calude [DKN12]. Calvia [Luo10, Luo08]. Cambodia [HV06]. Cambridge [And94, And96, AKR^{+}09, BGF06, BMS04b, BC96b, BC96c, CC95, CKO01, Chr99, Chr00a, Chr01, CCMR02, CCMR03, CCMR05, CCMR06, CCMR07, CCMR09, CMMR11, CCMS11, CMSA12, CS02b, CDL12, CM04b, DG06a, Den85, DKR02, FV01, FV03, FR01a, FFRS05, Fis08, Fit84, Fre96, Gae96, Gol96, GCM08, GJRS12, HI93, HRC06, Hug91, JKRR04, JBJ94, JQ04, KG12, Kil05, Kit88, KR01, LYCV08, LNS11, LKT08, Lom97, LST^{+}12, LS10b, MMK^{+}05, Nao04, PRJ95, PJ12, PZH96, RV02, SLF91, SMCM07, TC99, WC98, WMdLA10]. Camera [IS12c]. Camera-Based [IS12c]. Caminha [BDPS02]. Camogli [MMP07]. Campiglio [BCL03, MSW05]. Campinas [LM98c, WC95]. Campo [Bar08a, dRCVT11, DZGS06, dSK12]. CAMS [MTH05, MTH06a, MTH06b, MTH07a, MTH07b, MHD09]. Canada [AGB^{+}11, AMW07, AW01, A^{+}08, Akl95, ACD^{+}07, AN95, Ama08b, ASR^{+}07, AMRŚ08, ALPH10, AFG^{+}11, AHSC11, AKS09, BBD^{+}10, BS11a, BMR^{+}04, Ber08, BY07, BGS11, BK02, BFSN10, BDK10b, BJ12a, BG07a, BM02b, BGJ11, BEM06, BAP^{+}05, BFvdM10, BRP09, BC98, BHM^{+}02, BTF05, BKO04, BFKO06, BL11, CAD^{+}02, CYM^{+}11, CCS04, CKP^{+}12, CBG^{+}10, CP06a, CDCW09, CKK^{+}09, CM09a, CFG96, CS02a, CDN12, CFP94, CMQ96, DM01a, DFPV11, DSS89, DFK91, DSS91, Deh93, Deh97, Deh99, DSS03, DLOS05, DSZ07, DGST09, DF11, DMRS07, DOSY05, DS11c, Dwy01, EP03a, EP03b, ELN^{+}10, EW03, FK10, FKB02, FMP11a, FMP11b, FMP11c, FHH11, FGK^{+}10, FGM92, FGL96, Fri04, GJ09, GLSY10, GAL12, GFV00, GS00a, GK01, GP08b, GN08, GV09, GS94, GDG11, Ham00, HH04, Har01, HJ12, HA00]. Canada [HM05a, HYMZ11, HV98, HNPP07, IS11b, JPBT05, Jac04, JRSN09, JP07a, JSK10, JP11a, JP04b, JY14, JS94, KC05, KC07, KC09, KKGK11, KC11a, KC11b, KO91, KV04, KL05, Ker01, KD05, Kir93, KW07b, KH02, KI12, Kro00, KKGT03c, KKGT03a, KKGT03b, KR06, LLT07, LM06a, LV02, Lin07, Lin12, LMCH12, LN01b, LSIT11, LOH05, MZ07a, MWD06, MDHB11, MKPV05, MS05, MDY07, MMDN12, MZ04, MRP11, McC96, MO08, MH04, MN98, MCSN10, MR07b, Mil12, MMS93, MV12a, MNS06, MMT08, MG07b, MB98, MGSM12, NP12b, NOOS05, NBK04, NW10, NH03, OSV11, OYA04, OK96, PSS^{+}10a, PU04, PMWO02, PG01, PBK03, PT06a, RM97a, RWY97, RS12, SN08, SMB03, SS01, Sch09b, SDH05, SW08, STS04, STA^{+}11, SYP^{+}05, Smi12, ST07, ST01, STY^{+}07, SM01, SRL12a, SRL12b]. Canada [SRL12c, SWSY11, Tah00, Tan11, TM99, TG04, TBKO10, THX^{+}07, Tom92, TDK^{+}10, VY01, Vos03, Wal04a, WGMP09, WS96, WvdH03, WG11, Whi98, WY98, WMT96, WWC02, WANL08, XCd03, YDW08, YLW^{+}07, YSP^{+}10, YRWS11, YSL12, YP01, ZSSM03, ZEL04, ZY01, vBC08, vdPS08, vP93, Lau93a]. Canada-France [CFP94, GAL12]. Canadian [Ber08, BL11, CFG96, CS02a, FK10, GJ09, GS94, Ham00, KL05, KW07b, KI12, LM06a, McC96, MN98, SM01, TG04, XCd03].

Canaria [AAB^{+}06, ANF08, CAS^{+}09b, CH99, DPA05, DPA07, LMB08, MDP04, MDBF01, MDPQA09, MDPQA12a, MDPQA12b, PMD97, TGAH11, VSH11, vdL98, vdL01]. Canary [BHHK08, FSdlPT11a, FSdlPT11b, MMMD97, MÁ05a, MÁ05b, Sio10, SCD^{+}10]. Canberra [BvLR04, GN98, HSVZ10, McK99, MS02a, MS03c, PHH10, ZL96]. Cancer [MDY^{+}10, MDHB11]. Cancun [FMC04, FR00, MTOK06, MTCOBH11, MCA^{+}07, Raj02, Rol00, SC05]. Canne [COBV11, CDFP11]. Cannes [AJ01a, AJ01b, Ber00, MMM01]. CANPC [FL00, PS98a]. CANPC’97 [PS97a]. CANPC’99 [SL99a]. CANS [BLO^{+}07, DWML05, FHW08, GMO09, HWG10, LTW11, PMC06]. Canterbury [BDS04, CFB^{+}05, CZ05, CPR12, DC12]. Capabilities [Bee02]. Capacity [SF10]. Caparica [ABFR07]. Cape [ABGL09]. Capital [vdBK05]. Capri [BGR08, CAdB01, Mio90]. CAPTECH’98 [MTT98]. Capture [ERK07, MTT98]. Card [DFPS06, GS08, Pro11, AJ01b, GLIC10, QS00]. Cardiac [CPR^{+}10]. Cardiff [ABFK10, Emb98, GJS08, HRM^{+}10, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, WT07]. cardiology [Sac04]. CARDIS [DFPS06, GLIC10, GS08, Pro11]. CARDIS’98 [QS00]. Cards [AJ01b, SMBQ07, AJ01a]. CARE [GDG11, Hol07, Ria08, Ria09, RtTMP10, RtTMP11, RtTM12]. Carlos [dVGR04]. Carlsbad [CJK88]. Carlucci [SS06e]. Carmen [WB09]. Carnegie [BRW85, KC84]. Carnegie-Mellon [BRW85]. Carolina [Der89, RZ94, RS97b]. Carolyn [ADM11]. Cartagena [CBGCCZ12, PDMFF12]. Casablanca [Vau08]. CASC [GMV06, GMV07, GMV09a, GKMV10, GKMV11, GKMV12, DFT04, GMV05]. Cascais [Hof01, QBA01]. CASDMKM [SXC04]. Case [AW01, ABB99, ABMH08, AB03, BM10, CP02, KMK10, MS12b, MW09, MAR05b, RW11, RBGG06, WR07, Ber99, BB96, BP01, BMS96, FG04b, HKM95, LL95, RR89, SF96, SC98, Wat95, WSL93, WAR94b, EW91, AW12, AW01, ABB99, AB03, BM10, LP97, LBSBW98, MW09, MAR05b, RW11, VA95, WR07]. Case-Based [AW01, ABB99, ABMH08, AB03, BM10, CP02, MW09, MAR05b, RW11, RBGG06, WR07, Ber99, BP01, FG04b, HKM95, SF96, SC98, Wat95, WAR94b, AW12, BM10, LBSBW98, MW09, MAR05b, RW11, WR07, AW01, ABB99, AB03, LP97, VA95]. Caserta [EEM^{+}11]. CASL [BM03b, Mos04]. CASSIS [BBH^{+}05, BGH06a, MR95]. CAST’94 [OK96]. Castel [ETG02]. Castellon [MdPA98, dPMA98]. Castiglioncello [CPP07, PZ07b]. Castillo [Bar08a]. Castle [AFKL04, AHK07, AJ94, AKR03, BdBC01, Bar08a, BRS^{+}07, BHGP02, Ber03, BMS95, BM02a, BJS93, CBN99, CRYS09, CMOZM11, CKLM09, DGT96, DFP^{+}12, Die02, DFN97, DS06, ES02, FS05, FS99a, FW92, FLMS05, FLMK06, FLT07, FKLT08, GKS^{+}07, GKT09, GKL^{+}11, GSS00, Hag02, HHLR08, HRRY08, HS98c, IPSK04, Jan89, Jan92, JLST06, JLM00, KSHS01, KN91, KEM07, KHG01, KS05b, KS04c, Kra99, LBCO04, LS05, MBS05b, Nag90, RSS06, RBS93, RHD08, RME04, SST00, SKP07, SE94, STT04, Win04, ZS05b, GT05]. Castro [IT06, MSX10]. catalogue [US90]. Catalonia [LP00, MOFM10, OP97, PB97, SPSV^{+}04, TN04]. Catalonian [ETG02]. Catania [ABLM05, ABM02, BdR94, MTS^{+}03, MT^{+}03, NCBT04]. Categorial [MR12b]. Categorical [Ehr89]. Category [PPR87, PHSZ06, Man75, MR97b, Pit86, Pit89, Pit91, PRJ95]. Category-Level [PHSZ06]. Catharines [MWD06]. Catone [JS08]. Cauchy [Naz94]. CAV [AH96, BJ06, BM09a, DH07a, ER05, GQ11, GM08b, KP15, MS12a, TCJ10, AP04b, BL02b, CBF01, Dil94, ES00a, HS03, LS92, Wol95, vP93]. CAV’93 [Cou93a]. CAV’97 [Gru97]. CAV’98 [HV98]. CAV’99 [HP99a]. Cavtat [BHM^{+}04, DM06, LGTB03, LGBT03]. Cavtat-Dubrovnik [BHM^{+}04, LGTB03, LGBT03]. Caylus [Aus93]. Cayman [Syv02]. CBDAR [IS12c]. CBP [tHBP08]. CBSE [CSR08, GHC^{+}06, GRP10, HCS^{+}05, LPH09, SCHS07, CSSW04]. CC [Bod05, Gup10, Ham91, Hen08, Kno11, KO07b, MZ06b, O’B12, dMS09, Due04, Fri94, Hed03, Hor02, Kos98, Wat00, Wil01a]. CC’92 [KP92]. CC’96 [Gyi96]. CC’99 [Jäh99]. CCA [BBH01]. CCBR [SLCT11]. CCHSC [Ham89]. CCIA [ETG02]. CCIPL [DO89b]. CCITT [BLW82]. CCIW [STT11, TST09]. CCL [Jou94]. CCL’99 [CMT01]. CCMMS [JSS08]. CCPI [ADB^{+}12a, ADB^{+}12b, GVT^{+}11]. CCPSD [AM91a]. CCS [Fei03, San02, Tau89]. CCS-8 [San02]. CCS-9 [Fei03]. CD [DE05, QBY^{+}12, Bis02, EW04]. CD-ARES [QBY^{+}12]. CDB [Gae96, KR04b]. CdeA [Pau03]. CDMA [LhK03]. CDS [CMSM^{+}10, MGSM12]. CDVE [Luo06, Luo07, Luo08, Luo09, Luo10, Luo11, Luo12, MDH10c, Luo04, Luo05]. CEE [HKM^{+}11, MNW08, SSZ12]. CEE-SET [HKM^{+}11, MNW08, SSZ12]. CEEMAS [BLVV07, DKN02, MMP03, PVP05]. CEFP [Hor06, HPSZ08, HPZ10, ZHP12]. Celebration [HHG11]. Cell [PPdV12, LL95]. Cells [LST^{+}12]. Cellular [BCT02a, BMUV10, ECB06, SB12a, SCH04c, UMN^{+}08, WLS89, Sip97, BCT02a, BMUV10, ECB06, SB12a, SCH04c, UMN^{+}08]. CENEX [DFF06]. CENEX-SDC [DFF06]. Centenary [CDL12]. Center [BZ83, GHR89]. Centered [FP08, HKW09, HKW^{+}10, HKW11a, HKW^{+}11b, HKW12a, HKW^{+}12b, KSFN08, Kur09, Kur11, Lam00, KEM07, SS04, WFB12]. Centers [HdMKS12, Sch74]. Central [BLVV07, DKN02, HKM^{+}11, MMP03, MNW08, PVP05, SSZ12, BRR87b, Hor06, HPSZ08, HPZ10, ZHP12]. Centred [BFGL10]. Century [KM11, RLKHL12]. CEPES [MSZ90]. CEPES-UNESCO [MSZ90]. Cernobbio [MBC^{+}97]. Certified [HM12b, JS11b]. CESC [CPR^{+}10]. Cesena [CH96]. Ceské [MM03]. Ceský [Kuc̆02, KK07]. Cesme [HWW^{+}10]. CFLP [EKLM87]. CG [CM09b, HBN06, NV08, Tan11, TD07, vdHCD07, vdHXMW08, vdHIP11, BN91a, MF01, SMB03]. CG’98 [vdHI99]. CGGA [ABKT11]. CGI [NPS06]. CGWS [ADB^{+}12a, ADB^{+}12b]. Chains [Her02]. Chair [DFS08, DFM11, DFT12]. Challenge [CPR^{+}10, KKW11, KKW12, PBCF08]. Challenges [CKP^{+}12, CB10, FSdlPT11b, HHG11, LSK08, MSWZ06, MS10d, MG07b, OLL10, PPCS12, PPPM10, QCDM06, SNY06, SPC^{+}08, SHRM06, TDK^{+}10, VU09, Vet11, WBHR08, YC11, DTV02, MCA08, Thi00a]. Challenging [HP09]. Chambery [FO95, Ral94]. Champaign [BHL^{+}03]. chance [Had94]. Changchun [ZZP^{+}10]. Change [Che99a, CKH10, WPRW06, Fre98, FM91, MSS12]. Changing [AGT98, HTMC09, TOL08]. Changsha [AS11b, CC09a, CWZ11b, HZBF10, HZGZ10, PWY08, TNYL11, WCO05a, WCO05b, WCO05c, WJ05]. Chania [AB12, CCPP12, LDC^{+}09, PR11]. Channel [SH12b]. Chantilly [BKH^{+}09, HKM95, PS09b]. Chaochen [JLW07a]. Chapel [CPC^{+}99, Der89]. chapters [MT99c]. Characterization [NP11a, NP12a]. Charges [SCKR03]. Charging [CMS11, RST09, SPSV^{+}04, SCKR03, SRT06, SSKR02]. Charleroi [Van05]. Charles [LLM09]. Charleston [EN06, Hen06, REPH06]. Charlotte [Mic90, RZ91, RZ94, RS97b, RO00a]. Charlottesville [PNB04]. Charme [BP05, CE95, GT03, MM01, MP93]. CHARME’99 [LK99]. Charting [MP06]. Chateau [BCM00]. Chatillon [Pod95]. Chatillon-sur-Seine [Pod95]. Checking [BH10, BE07, CP07, Cou05, DP12, EL07, EN06, GM04b, GM11b, GV08, HMP08, JS11a, JMO09, KR12, LPZ08, Păs09, PW09, SL03, Val06, vdPW10, vdMS11, BR03, BL02a, Cor02, Dam99b, Dwy01, God05a, HPV00, Sud95, Zuc03, Qui04]. Cheju [Cho02a, Cho02b, Kah03, Kim01a]. Chemical [NM08]. Chemistry [PS07a, PS08a, DMW96]. Chemnitz [NR98]. Chengde [LMC12]. Chengdu [CEQ^{+}11, DTW03, GKC^{+}12, LWDM12, LNW^{+}12, TWQ^{+}12, TLZ^{+}08, WYZ06a, WYZ^{+}06b, WLGB^{+}08, W^{+}12, YZC^{+}07, YYS^{+}12]. Chennai [AR99, AR98, BC11a, CV04b, CFRR10, LM04, PDSD10, RRR99, RS10, RS09a, RD01, Roy05, SRY07]. Cherbourg [ELNM08]. Cherbourg-Octeville [ELNM08]. CHES [CG09, GM06b, MS10a, OR08, PV07, PT11, PS12c, RS05, JQ04, KKP02, KP99, KP01b, KNP01, WKP03]. Chester [Erl06, FG06]. Chi [NHJ12a, NHJ12b]. Chiang [Mah04, TKSN11]. Chicago [ASP01, BCB99, CPFD^{+}03a, CPFD^{+}03b, CGKM11, DL06b, ES00a, EG10, GRB^{+}07, GHSW04, HM03, HMBD09, JP03, Lom93, Mar93, MF06, MAR05b, NZS^{+}06, SLLS03, WW02]. Chile [BYGP95, CHK06, GLMP07, HP02a, HN12, MR07a, RMK07, SK11, VRHJ09a, VRHJ09b, ZBY07]. CHILL [BLW82]. Chilton [Rea01]. China [APR^{+}06, ADDL08, ACL12, ACK^{+}07, ABKT11, AGL^{+}07, AS11b, ALZ11, ACL^{+}04, Ayk07a, Ayk07b, BBCL08, BW03a, BLW09, BYLJ11, BW11, BGG^{+}12, BRL08, BTZ^{+}10, BDDM12, BC11b, BCF^{+}02, BZ08b, CCL06, yCCZ07, CLKL09, CHKL10, CIW06, CPR^{+}10, CYGL04, CNX05, CSJD06, CLWC08, CFZ10, CZF10, CHB^{+}12, CT06, CPSA10, CWC^{+}07, CZG^{+}08, CWZ^{+}11a, CLL^{+}09, CGL^{+}11, CCM^{+}04, CPGL05, CDLZ06, CPZ07, CJ08b, CLZ^{+}09, CC09a, Che10, CY10, CTYG10, CTS10, CWZ11b, CY11a, CFLR12, CYZ12, CP06b, CLZ06, CEQ^{+}11, CWL01, CFK^{+}12, CL04, CvSW04, CY12, CP98b, CP98c, CBS^{+}11, CCAS12, Dai07, DW07a, DD06, DY06b, DWWL09, DHX09, DMLW11a, DMLW11b, DMLW11c, DWML05, DSZ10, DTW03, DLW^{+}07, DZ10b, DXZ07, DZ94, DL95b, DHP09, DFW^{+}11, Duf07, FWY05, FLY06, FWZH06, FT04a, FX08, FCLL02, FKW08, FHW08, GWY99, GML^{+}08a, GLW^{+}12, GKC^{+}12, GM03, GLW07, GPYD09, GHSY06, GLC^{+}11a, GLC^{+}11b]. China [GZ06, GSV05, GPL99, HOQ97, HTW02, H^{+}05, Har07, HHP09, HLKX12, HPR05, HhKRT06, HNS^{+}12, HL91, HW08a, HLCZ11, HSSV12, HM12c, HZH05, HLI06a, HLI06b, HLI06c, HHL07b, HHL07a, HWLJ08b, HWLJ08a, HYP^{+}09, HZBF10, HZGZ10, HCS11a, HCS11b, HGBF12, HGGG12, HGPH12, HJBF12, HMJG12, HL00, cHPkC^{+}07, IMY^{+}07, IS98, IAL^{+}07, IHX^{+}08, IS12c, JZR09, Jac07c, Jac07d, Jac07b, Jac07a, JYZ04, JE06, JWH06, JL97, JNPV10a, JNPV10b, JNPV10c, JDAW12, JWbG^{+}06, JWG^{+}06, JPXS04a, JPXS04b, JGXC04, JRJ05, JRPP07, JLW07b, KLZ07a, KLZ07b, KCYL08, KY07a, KWCW06a, KWCW06b, KWCW06c, KYT01, KFfKL11, LC06, LYL11, LZL11, LLW06, LLCL05, LCY10, LWDM12, Leo00, LWM^{+}08, LPC^{+}11, LLT^{+}04, LWF04, LSDN04a, LSDN04b, LST^{+}05, LWD05, LOS05, LBO05, LZhL06, LFIM07, LLIH07, LJJG07, LZS^{+}08, LLM^{+}08, LFP^{+}09, LFJI10a, LFJI10b, LJS^{+}10, LCB12, LNW^{+}12, LEP^{+}10, LTW11, LYL06, LYhL^{+}01, LCY03, LSL04, LA05, LH06, LL06]. China [LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LWYN09, LR09, LLWL09, LDXZ10a, LDXZ10b, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, LC11, LMC12, LLTY12, LZ00, LZ05, LSU07, LZK11a, LZK11b, LZK11c, LM08b, Luo07, LSW^{+}10, Luo11, LCY^{+}11, MJYT06, MRN07, MCA08, MDY^{+}10, MXZ05, Mei08, MSW02, MLTC11, MSG06, MZWZ05, NPS06, Nob12, Oho03, OiDP98, OW07, PCG^{+}05a, PCH^{+}06, PAD^{+}06, PZE^{+}08, PCK^{+}12, PKM^{+}10, PZ08, PFCS05, PSPH^{+}10a, PSPH^{+}10b, PYLW08, PZ09, PMC06, PF07a, PWY08, Qin10, QOZ01, QGZ03, QMLW05, QIW07, QMW09, QSWL11, QLXL07, QLK^{+}10, QLK^{+}11, RWA06, RSW12b, SM11a, SR07, Sch07, SLH05, SLZH07, SLB05, SLL^{+}06, SLL^{+}07, SPÖ^{+}10, SWJX12, SXC04, SS06b, SvADS07a, SvADS07b, SvADS07c, SvADS07d, SKK08, SLC01, Shu07, SS07c, SS07b, SLSW12, Ste07d, Ste07b, Ste07c, SZ11, STLK05, SZT^{+}08a, SZT^{+}08b, SLCT11, TWQ^{+}12, TSG00, TLF01, TST10a, TST10b, TSCW11a, TSCW11b, TSJ12a, TSJ12b]. China [TWYhL01, TLZ^{+}08, TKCW11a, TKCW11b, THL11, TTP05, TYZ11, TNYL11, TCLH10, Ued10, VHM07, WYL01, Wan01, Wan03, WTZ^{+}04, WCO05a, WCO05b, WCO05c, WLY05a, WLY05b, WLY05c, Wan05b, WJ05, WYZ06a, WYZ^{+}06b, WJS^{+}06, WPSY06, WLC^{+}06, WPRW06, WmCL07, WLGB^{+}08, WFZL09, WDGL10a, WDGL10b, WYSS10, WGLL10, WZD11, WLO^{+}11, WS12, WYP12a, WYP12b, WJL^{+}12, WZJX12, W^{+}12, WZH^{+}12, WZH^{+}07, WCGB05, WYX08, X^{+}12a, X^{+}12b, XYM^{+}07, XBS^{+}10, XCR11, XHXL08, XLHX08, XZCL07, XYZU11, XBM12, YJ04, YZZ^{+}05, YJMU06, YJS^{+}06, YW06, YZC^{+}07, YYS^{+}12, YLWY06, YKF^{+}06, YX04, YWG04a, YWG04b, YC11, YLLZ04, YKL06, YHZ09a, YHZ09b, YHZ09c, YGL^{+}10, YLC^{+}10, YKU11a, YKU11b, YZCW07, YLL09, ZTTY12, ZPT^{+}06, ZJ04a, ZHF04, ZLJ05, ZTY^{+}05, ZJ06, ZOCJ07, Zha07, ZYBX08, ZLK10a, ZLK10b, ZZP^{+}10, ZCDT10, ZS10, ZHLW12, ZZZL12, ZGT05, ZJL06, ZZ99, ZLY^{+}07a, ZLLC09, ZCGH11, ZYH03, ZJXC03, ZZO03, ZOM05, ZLS^{+}06]. China [ZLY07b, ZW09, ZY10, ZZLL10, ZYRH06, ZF05, ZYW08, vDLJ11, vHV08, vdHXMW08]. China-Japan [ACK^{+}07]. Chinese [DEM96, DJ08, LLT^{+}04, LSU07, SXC04, SLCT11, HMCL06]. Chinese-German [LSU07]. Chios [RG03]. Chisinau [MR01, GMV06]. choice [Chr91]. Chongqing [BTZ^{+}10, CFZ10, CZF10, HHP09, TSCW11a, TSCW11b, Wan03, WLY05a, WLY05b, WLY05c]. Chongquing [WPSY06]. CHOROCHRONOS [KSF^{+}03]. Christ [CSN14]. Christchurch [ET01a]. Christoph [SW10]. Chronic [DPNM12]. Chu [CCL02]. Church [CSN14, EK09]. CIA [KRP06, KHPS07, KPP08, KK97, KW98, KK00, KZ01, KOS02, KOOL03, KOK04]. CIA’99 [KSW99]. CIAA [BMCCM11, DOSY05, DS11c, FLS06, HŽ07, IY06, IR08, Man09, MR12c, CM03a, ID03, WW03, YP01]. CIAC [CFI06, CD10, BBD97, BGP00, BCP94, PPS03]. CIAO [MTH06a, MTH06b]. CIARP [AMGJ12, BCE09, BC10b, MTOK06, RMK07, RSK08, SK11, STC04, SL05, SRS03]. CIBB [BVAT12, MTV09, MPT10, RL11a]. CIC [LhK03]. CICLING [Gel05, Gel06, Gel07, Gel08, Gel09, Gel10, Gel11b, Gel11a, Gel12a, Gel12b, Gel01, Gel03, Gel02, Gel04]. CICM [CDCW09, JCC^{+}12]. CICW [LCY^{+}11]. CIE [CLT05, ASLM09, BBLT06, BDL08, CLS07b, CDL12, FLMG10, LNSS11]. CIKM [FNY93]. CIM [ALM95]. Cimino [CW94a]. Cincinnati [BD08]. CIP [BBB^{+}85, BEH^{+}85, BW91]. CIP-L [BBB^{+}85]. CIP-project [BW91]. CIP-S [BEH^{+}85]. Cipher [RB08]. ciphers [DXS91]. Circuit [AGCP^{+}11, AS07c, CM03b, HAB02b, MPK04, MvL10, SPB00, SM09, VAM06, vLS11, KK95, PVV05]. circuits [Klo87]. Cirencester [Boy95, Dar97, Gal07, Hon01, Par09, Pat03, Sma05, Wal99a]. CIS [H^{+}05, WmCL07, ZHF04]. CISC [FLY06]. CISE [CBS^{+}11]. CISIM [CCSW12]. CISIS [HC11]. CISMOD [Bha95]. CIT [DG04]. Cities [vdBK05, II00, TvdBI02]. City [BDK10b, BI04, CH04, CMB^{+}08, FRL07, FKM08, Gel01, Gel02, Gel03, Gel05, Gel06, Gel07, Gel09, Heb95, HK01, JRL09, Kro00, LLS00, LM06a, LMT00, LMA08, MA11, Mil12, MMS93, MKB^{+}09, MAFSS04, NLŚ10a, NLŚ10b, NHJ12a, NHJ12b, RCLE05, Rus00, SS07a, SKPG12, SBG04, STWS07, WVDS05, YGGJ04, Zho01, ZRTS03, ZJS^{+}02, tTVH^{+}12]. Cividale [MSH08]. CIVR [BHL^{+}03, EKO^{+}04, LLC^{+}05, SNSR06, LSE02]. CJCDGCGT [ACK^{+}07]. CL [Llo00]. CLA [YNB08]. Clara [Ban92]. Classes [Han81]. Classical [CS00a, ACG^{+}07, BJK06]. Classification [GJTS06, Mül07a, Mül07b, QCDM06, SG07]. Classifier [BBMB^{+}08, BBD^{+}10, BKR09, EKR10, HKR07, KLT^{+}07, LSW03, RK02, RKW04, SKR11, KR00b, KR01, LSW00, LSW01, LSW02, OPKR05, WR03]. Clauses [HR99b]. CLEAR [SG07, SBF08]. CLEF [AFP^{+}10, CFH^{+}12, FGK^{+}11, Pet01, PBGK02, PGBK04, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, PBGK03]. Clermont [CP12, KS07c, RMTW11]. Clermont-Ferrand [CP12, KS07c, RMTW11]. Cleveland [CU08]. client [Sch93a]. client/server [Sch93a]. CLIMA [DL04b, DFN10, DLGJ10, FST09, FvdTDG12, IST07, LT05, LTÅ^{+}11, SS08a, TT06b]. Clinical [CMSM^{+}10, MGSM12, YC11, YSL12, YHV12]. Cloister [BFKO05]. Clos [DGG^{+}10]. Closed [Vel94b]. Cloud [DPZ12, HHMT12, JZR09]. CloudCom [JZR09]. Clouds [AR10, VAR12]. CLU [Lis81]. Cluj [HPSZ08]. Cluj-Napoca [HPSZ08]. Cluster [GNTF02, JM02].

Clustering [BHVV09]. clusters [HR99a]. Clustweb [LMT^{+}04]. CMBSB [NSM07]. CMC [GHP^{+}11, GPR^{+}12]. CMC’98 [BB01b]. CMCS [PS12a]. CMLSA [HRK^{+}07, SPC^{+}08, TDK^{+}10]. CMMR [KMVY06, KMYJ08, Wii05a, YKMJ09, YAKMJ10, YAKMJ11, YAKM^{+}12, Wii04]. CMMR/ICAD [YAKMJ10]. CMS [CVLL12, DSB10, DLNU11, DC12, DKU05, LM06b, TDK^{+}10, LM03]. CMSB [CG07, DS05a, DG09a, HU08, Pri06, Pri03a]. CNL [Fuc10, KF12, RF12]. CO [CS05b, Gho10, SS09b, vdBGG10, DO02, GW96, Jon00, EM02, FR99b, HS03, Lee01, Bui87, Oss99]. co-located [Jon00]. Co-oP [Bui87]. Co-ordination [Oss99]. Co-sponsored [DO02, GW96]. Coalgebra [CKC11, KLT09, MMH07, FHRR05]. coalgebraic [BCG02, PS12a]. Coast [Boy01, CKH10, HNF08, OT07, Pap95, RAW07, SZ93, WLP^{+}09]. CoBuild [SKB98, Str99]. COCOA [DXZ07, DHP09, Lin12, WZD11, WD10a, WD10b, YDW08]. COCOON [CL06b, FD11, GMV12, HW08a, Lin07, Ngo09, TS10, Wan01, Wan05b, Asa99, yCW96, CM04a, Du00, HK98b, IZ02, JL97, WZ03]. COCOON’95 [DL95b]. COCOS [JNS05b, BJN03]. Code [Rüt98, Ste05]. Codes [BS01b, BL07b, BAH09, FILP06, BC88, Cal86, CMM93, CGM95, Fos99, FHP03, Gur04, HP89, MMR91, Mor89a, MM97, Pol86, Sak91, Chi79]. Coding [Che11, Gal07, Par09, Pat03, Sma05, Wal99a, Ytr06, Boy95, CC91, Coh92, Coh94, Dar97, Hon01, Bar08a, CLL^{+}09, CGL^{+}11, CG88, CW89]. Coevolutionary [Rey04]. Coference [ABS01]. CoFI [CR01a]. Cognition [BKLM07, FHW98, FBHW00, FNGW08, HSB^{+}10, MÁ05b, SR07, SEG09, SF11, SSU12, CM95, Fre03, FKKB^{+}05, SSW92, BFJV97]. Cognitive [CF10, EHMM09, GTWC09, Har07, Har09, Har11a, MÁ07a, OSR^{+}02, PR07, PT09, PBCF08, PBSB09, WPE05, ZHLW12, And95, FM99, Lem97, SW75, vdV84, BND01, CV08, CN06, LSU07]. coherence [Mac06]. coherent [HR99a]. Coimbra [BMO02, BPS06, BMM07, CVTP12, CC97b, De 10b, GK09, KOL^{+}04, MZCC09, RCB^{+}04, SKB10, Smi01]. COIN [CvRVSN12, DFPV11, HMBD09, NVSB^{+}07, PAV^{+}10, SPON08]. CoLIS [CR05b]. Collaboration [CCM^{+}04, GP07, GP09a, GP11, Gol11b, IFV07, BDGM05, Smi98, CKOS09, HHJZ12, KHL10, Pad99, VGB11]. Collaborative [Han02, Kam99, CN98, GDG11, NM08]. Collection [PAC08, FP90]. Collective [JNH11a, JNH11b, NKC09, NK10a, NK10b, Ngu11a, Ngu11b, Ngu11c, NHJ12a, NHJ12b, Ngu12a, Ngu12b, Ngu12c, PCN10a, PCN10b, PCN10c, DTF98]. College [BKA11, BR12b, DJ08, FHM82, PT06b, Rau04a, SYNC11, Sha86, TDS94, YGE12]. Collins [FR99b]. Colloqium [EKLM87]. Colloque [Rob74, Rob80]. colloquim [ADCR89, CIM^{+}05]. Colloquium [AM91b, AM91a, AGK^{+}10, ADG^{+}08b, AHS11a, AHS11b, AFvZ02, AM03, AMSM^{+}09b, ABdJS10, AKR^{+}12, Arn90, BM03a, BCC06, Bra07, BPSW06a, BPSW06b, CDGW10, CP11, CCM08, CGvR02, CMPW12a, CMPW12b, DN88, DBM09, DO89a, DO89b, DKLS04, Ehr85a, Ehr85b, EEK^{+}87, EH12, FHY08, FG06, FZ86, GLM93, GLM97, HDBM11, JLW07b, KY11, KŽ10b, LS88, LM09, Loe74, MD96, Nie09, Par84, PSE10, PR05, PZ07b, Rao92, RKM^{+}12, RD12b, SKS^{+}06, SG10, SF08b, WMR00, WTM^{+}02, AS94, ADG^{+}08a, AMSM^{+}09a, ACJT07, AB81, AB78, AP83, BLPW03, Bra85, CO94, CG88, CW89, DGMS97, DCM82, DR81, Dia83, EK81b, FG95, HS98b, HW05, Kir96, Kot86, KS04c, Kui92, LSW98, LKC93, LA05, Mau79, MM96, MAA91, NS82, Oli00, OSvL01, Ott87, PS04a, Pat90, PR84, SS77]. colloquium [SKMR94, Tis94, vWN99, vLdB80]. Colmar [Lor01a, Lor01b]. Colocated [PS12a]. COLOG [DB04, MLM90]. COLOG-88 [MLM90]. Cologne [Jou09, WKP03]. COLOGNET [HCF06, AFR^{+}04, FPFR05, O’S03b]. Colombia [CBGCCZ12, PDMFF12]. Colombo [MVCM10]. Colonial [Mar01]. Colonoscopy [YC11]. Colony [DGB^{+}06, DBB^{+}08b, DBB^{+}04]. Color [STT11, TST09]. Colorado [CH98]. COLT [AM05, BG07b, LS06c, HW01, KS02d, SW03a, STS04]. COLT/Kernel [SW03a]. Columbus [HSB^{+}96, XKGS12]. Columnia [CMQ96]. CoM [DMB^{+}11]. COMBEK [MTH08, MHD09]. Combinatoral [GW11b]. Combinatorial [AB11, AJRS03, BS06a, BJP12, BN04, CJRR07, CS02b, CvH07, CC09b, CM10, ES10, FW07, GJRR08, GJRR11, GR04, GR06a, GJRS12, HM12a, JK00, JLV02, JKRR04, JN01, LPR08, LMT10, LOH05, MH11, PT08b, RR04a, SSJR10, VW07, vHC08, vHH09, AG01, BC95, BM05b, BBRM98, CJRT05, CBW99, CMQ96, DJRZ06, DJNR09, FP96, GJRT01, HJRS99, JR98, JK05b, RG05, ABB^{+}11, ABK05, AL01, AP10, Apo92, Apo93, AH97a, AT02b, ACP05, BYCC03, BBA12, BBH08, CG94, DXZ07, DHP09, Erl06, FC98, FL08, FKM09, GU95, GS00a, GM11a, HM96, IS11a, IS11b, JP07a, JRR03, KS12a, KŽ04, KU09, LV06, Lin12, LOH05, MZ07a, MMMP12, PC99, REF^{+}06, SMD04, WZD11, WB09, WD10a, WD10b, YDW08]. combinatorical [BJK06]. Combinatorics [ACK^{+}07, CL06b, CM04a, FD11, GMV12, HW08a, IZ02, Lin07, Ngo09, TS10, Asa99, BC88, yCW96, DL95b, Du00, HK98b, JL97, Wan01, Wan05b, WZ03, ABC^{+}06, CPZ07, DEM96, Die90, PS97b]. Combinators [CCR86]. Combined [HNRW06]. Combining [AM91a, GS09a, KW07c, TSS11, Arm02, Gra05, KR00a]. COMINF [MTH06a, MTH06b]. command [Bee86]. commemorating [PSD^{+}07, PSM^{+}07].

Commerce [BMP00, BTQ03, FPRAW04, FS07a, FKB02, FMW01, LSJ06, LSgL^{+}06, LY01, NMP^{+}04, Per02a, Per08, AY96, BMP01, BTQ02, DS01b, DC01, FRA05a, KW98, KYT01, LM98a, MSY00, NS99, PSP^{+}02, BPW06, DB09, PW07, PW08, BBP04, FBSMS05, PBW05]. Commission [Omi00, PPF^{+}00]. Common [BM03b, KMK10, RRMP08, SDM05, Mos04]. Commonsense [PWZ98]. Commonwealth [PY05a]. Communicability [CFVV^{+}12, FdCLJ^{+}11]. Communicating [GH04, PVTG07, SHUH08, UDF06, HW03, KD05, Mil80, AJS05]. Communication [AAW04, ABDW10, AZ11, BW03a, BI03, BH07, BDZ08, BPZ09, BI11, BCP12, BPS06, BAP^{+}05, BFvdM10, CV04a, CM09a, CB06, CH01, Dig04, DvEF06, EKF12, ERRW03, EFZKM07, EVV^{+}11, EH12, FOGE^{+}09, FG06, GCC^{+}02b, Hug03, HHG11, Ish10, IS12a, IS12b, KM08b, KW10, KY11, KMMT12, KS04c, KT11, KŽ10b, LYCV08, MKR^{+}04, Mue06, NS81, NvRSS11, NBM09, PSE10, Paz03, PR05, PZ07b, ST02, SF08b, Smi05, SV12, VSTN06, WK08, dMLG03, vDKS11, BB95, BC04a, BB05b, BB01b, BDR99, BC05, Bol78, BS94, Bra99, BBB98, Chu03, CGP03, DN93, FL00, PS98a, PPF^{+}00, RF99a, SILS05, SL99a, SS06d, Ste94, Tuc01, VM99, YD98, vEHD05, Had96, PS97a, SFV^{+}11, VMB^{+}12, DG00]. Communication-Based [ERRW03]. Communications [BPS06, BCF06, CRW08, CY12, Cho02a, Cho02b, CK06c, CGGO03, COB06, DQBZ02, FLL03, GK06, GCC^{+}02b, HHLS08, LLMT04, Leh11, ML04, MBVTY11, MKR^{+}04, NdG04, NQL06, OKBMB10, Pal01, PCS^{+}07, QOZ01, QGZ03, QIW07, QMW09, QSWL11, SBV^{+}11, SQL10, SCKR03, SGD09, VBS^{+}10, vdBHOS09, vdMBD08, BMM07, BCKT05, Gun94, HOQ97, Hut96, LhK03, LQO04, Omi00, Pla96, PPF^{+}00, QMLW05, Wil99, YRMD05, DSB10, DLNU11, DC12, DKU05, Hug03, LM06b, LM03, Sin94]. Communicative [BMN05]. Communities [OZ09, OZ11, PKSS02, Sch07, Kro00, Ish99a]. Community [BRU06a, BHU03, BBB06, Ish99a]. Como [AH98, BFH07, BFBS11, Boe01, Boo91, CDSS07, FB09, Mil01]. Comoa [WTZ^{+}04]. CoMoGIS [ALS^{+}05, RBsC^{+}06, WTZ^{+}04]. CoMoL [HP09]. comparative [Hil78, BEM06, CM09b, FLT07, GKST11, Lag05, MH05, NV08, PGBK04, Tan11, TD07]. Comparing [RRMP08]. comparison [KS97]. COMPASS [BC93, EO94, HOD96, RTA95]. competition [MBC^{+}97, Lib07, Gur04]. Compilation [PA01, RSB07, Ham89, Har01, LO98]. compiler [Har77, Jon80, KHZ82, MO97, BE74, BE76, Bod05, Due04, Fri94, Gup10, Gyi96, Ham89, Ham91, Hed03, Hen08, Hor02, Jäh99, KP92, Kno11, Kos98, KO07b, MZ06b, O’B12, PA01, Pol81, Wat00, Wil01a, dMS09]. Compilers [AGP08, ACW07, Ama08b, ABRS06, CNmWH^{+}05, CMCS11, DKS^{+}07, Die03, Dwa00, GPCL10, Kra03, O’H98, PFD^{+}10, Rau04a, Sch04a, SEO^{+}09, Ste07a, SDK^{+}08, Ste09a, Ste11a, Ste11b, Ban92, Ban93, Ban94, CF00a, CPC^{+}99, DM01a, ELM05, FS03, Ham89, Ham91, HG96, HSB^{+}96, Li98, MMG^{+}01, MB98, Pin95, PT06b, Seh97]. Compiling [Pet99]. CompIMAGE [BBH^{+}10]. complementarity [Koj91]. Complementary [Don76]. Complete [Mos04]. Completeness [vM00]. Complex [BJRT06, CG12, CT02, CPSA10, FFM05, Geu98, LWZ09, NP11a, RTZ08, SKS^{+}09, ZSD03, ZZ04b, AFS89, AGT98, Atz93, BJM^{+}05, vM94, JMBS07]. Complexity [BGP00, CFI06, CD10, CPSA10, EH12, FG06, Gol11b, Hel08, HKP11, KY11, KS04c, KMR12, KŽ10b, MSS02, PPCS12, PSE10, PR05, PPS03, PZ07b, SF08b, Thi00b, TWSN07, BBH01, BBD97, BCP94, BHR84, De 87, Lei95, MP95b, MD01, Sel86, CKV08, Rüt98, Sch86, SS76, vM00, DK89]. CompLife [BGD^{+}05, BGF06]. CompNets [BBK12]. Component [AJRB09, AÖ12, BL12, BCT02b, DJAP07, KV04, PS03, PP04, RRMP08, REPH06, Bos01, BJ01d, CE00, GL05, ABGP05, Bis02, CPV03, CSR08, CSSW04, DE05, EW04, GHC^{+}06, GRP10, HCS^{+}05, LPH09, SCHS07]. component-based [Bos01, BJ01d, CE00, CPV03, CSR08, CSSW04, GHC^{+}06, GRP10, LPH09, SCHS07, ABGP05, HCS^{+}05]. Components [AdBB12, BV11, FN12b, Mor06, OSRS07, RSS06, Vit10, dBBGdR03, dBBG04, dBBGdR06, dBBGdR07, dBBGdR08, dBBM09, dBBHL10, AMU03, Gun94, dBBGdR05]. COMPOS [dRLP98]. Composition [AJ11, BW10a, BF09, CS05a, GDGB12, LS06b, LV07b, PT08a, GAN05, KS07a, Lin03]. Compositional [Geu98, Hoo91]. Compositionality [DHS10, Rao95, Zwi89, dRLP98]. Compostela [LFL05, MOBS06, MFÁ^{+}09a, MFÁ^{+}09b, MOTN07, RBTD04]. comprehensive [ALM95, Lin03]. compression [CW94b]. Compulog [Apt00]. Computability [ASLM09, BBLT06, BDL08, CLS07b, CDL12, FLMG10, LNSS11, CLT05, BBH01]. Computable [CSCR08]. Computation [AI06, AAA^{+}11, ADDL08, ACL12, ACD^{+}07, AP04a, ABGL09, Asa06, AiNOW11, BdQ11, BBB07b, BLSS11, BG12a, BDR10, BL06b, BGL09, BM02b, CCL06, CLKL09, CHKL10, CBC^{+}02, CIW06, CDP02, CDP^{+}06, CCD^{+}09, CHM^{+}10, CKPR11, CPFD^{+}03a, CPFD^{+}03b, CJRT05, CF09a, CC09a, CCP10a, CCP10b, CM09a, CI98, CLKK07, CvH07, CC09b, CM10, CVÉ07, DU09, DSM94, DdQ10, DW12, DPB^{+}04a, DPB^{+}04b, DCC^{+}10, DBD^{+}10a, DCC^{+}11, DBD^{+}11, DAC^{+}12, DJRZ06, DDI09, DFD04, DLJ12, ET01a, EV96, EM08, FSdlPT11a, FSdlPT11b, Fit84, FJM06, GVB12, Gol11b, GR04, GR06a, GN08, HM12a, HT06a, HdQ08, HNF08, IKO03, IM10, JKRR04, JWbG^{+}06, JWG^{+}06, KLZ07a, KCYL08, KK11b, KM08b, KLFK10, KCG09, LS12a, LT00, LdQ07, LN03b, LR05a, LNSS11, MT09c, MM08, MR12a, MH11, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b]. Computation [MSS02, iNR08, OT11, OdQ12, OKdQ09, OST11, PRG09, PRG10, PRG11, RF10, RiN12, RS10, RKS^{+}12c, Sar03, TW12, Tok07, Wal04b, tCZ07, vDKS11, vHC08, AR99, Asa96, Bos96, BGY96, BC04b, Bud85, BBL87, Bud91, yCCZ07, CC95, CCP96, CL96, CP98a, CDP^{+}05, CRL01, CC97a, CP99, CDG89, DD06, DSV01, DZ94, Esi93, Fit93, FT04a, Fre01, FU96, Gec81, Gia89, GMW79, HS98a, HM01, Iba92, Kah79, Kar77, Kar83, LIJ97, LK00, MP03a, Man75, MMR07, Mio90, Mio93, MCP93, MS95, MMMD97, MT05b, Ng79, Ng93, NE02, Niv98, RG05, Rei95, TJHS94, Tho99, Tiu00, TG02, Vää99, VvS99, Vol93, WCO05a, WCO05b, WCO05c, Waś96, WL03, Yao95, Zim92, ZVU99, Bor87, CLS07b, DKN12, KK11c, Neh99, Sch98b, Sko85, Ram96]. Computational [ABH12, AK03, AKT06, Ama08a, ASLM09, AGI^{+}06, BS11a, Bar02, Bat09, BCM08, BD12, BBLT06, Ber10, Ber08, BDZ08, BN91a, Bir12, BL04b, BvASD04a, BvASD04b, BvASD04c, BvASD04d, CSPC09, CRJ11a, CRJ11b, CPR^{+}10, CKP^{+}12, Cho12, CS02a, CMOZM11, DDHP05, DPC^{+}07, DWWL09, DMLW11a, DMLW11b, DMLW11c, DSS^{+}07, FMS10, GT08, GTW^{+}08, GT09a, GTM09a, GTWC09, GT09b, GT10, GTA10, GTM10a, GTM10b, GT11a, GTSS11, GT11b, GTM11, GTP12, GT12, GPP09, Ham00, HW01, Hof11, HKS09, IB07, Jha10, KS02a, KS02b, KUA00, KL05, KS02d, KW07b, KH09, LM06a, LMSS07, LWDM12, LFIM07, LLM09, LABM^{+}12, MRU12, McC96, MN98, MA03b, MS08a, MMK^{+}05, NK10a, NK10b, Ngu11a, Ngu11b, Ngu11c, Ngu12a, Ngu12b, Ngu12c, NRW04, Ong10, PAD^{+}08a, PDML11, PP11, PMGO05, Pri05]. Computational [PZ05, PCE06, PHPL06, PP06, PIMN06, Pri07, Pri08, PDAN08, PBP09, PBG^{+}10, PBPdV11, PPdV12, Raj09, SW07, Sas05, Sei07, SVA05, STDH02c, STDH02a, STDH02b, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, SH07, Ste12, SM01, TvdBI02, TG04, Thi00b, VW08, WDGL10a, WDGL10b, WAW01, XCd03, YC11, YSL12, YHV12, ZYW08, dA09, dSK12, dSTP11, vM00, AKU01, AKR03, BE95, BSBM05, CMT01, CLT05, DMR01, FM99, GM00, Gor03, GP89, Hel91, HR91, Jou94, LLP99, Lei95, LM96, Mei89, Moo01, PTRH05, PDHS03, Rei94, Ret97, Rob93, Tei93, ABKT11, ADJ^{+}01a, ADJ^{+}01b, AvAS06, AvASD06a, AvASD06b, AvASD06c, ANS^{+}09a, ANS^{+}09b, Alt01, BBH^{+}10, BD97a, BGD^{+}05, BGF06, BN91a, BVAT12, BHJ^{+}11, BvADS08c, BvADS08a, BvADS08b, CPS05, CG07, CVTP12, DS05a, DG09a, DL04b, DFN10, DLGJ10, DT07]. Computational [ÉEMS04, FFL^{+}12, FS99b, FST09, FvdTDG12, GS01, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, Gel01, Gel02, Gel03, Gel04, Gel05, Gel06, Gel07, Gel08, Gel09, Gel10, Gel11b, Gel11a, Gel12a, Gel12b, GFV09, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, GG07a, GG07b, GG07c, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, GLM93, GLM97, H^{+}05, HU08, HW01, HC11, HSSV12, HM12c, HLI06a, HLI06b, HKH10, IST07, IWC04, IKKU08, JNH11a, JNH11b, JN01, KS02d, KKGT03c, KKGT03a, KKGT03b, LGK^{+}04c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, LT05, LTÅ^{+}11, Llo00, MBT^{+}03, MRU12, MTV09, MPT10, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, NKC09, NHJ12a, NHJ12b, Nol88, PCN10a, PCN10b, PCN10c, PBK^{+}10, Pri03a, Pri06, Reu97, Reu99, Reu01, RL11a, Sac04, SS08a, SFS11, SÇS12, SPCG07]. Computational [STT11, SS89, SvADS07a, SvADS07b, SvADS07c, SvADS07d, SF08c, SvASD05a, SvASD05b, SvASD05c, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, TdLdOQ08, TT06b, TST09, V^{+}06, Vit95, WmCL07, ZHF04, ZYW08]. Computations [DLM07, MR01, Mar05a, CKPB07, DMW96, Klo94, MT86, Sta95]. compute [HR99a]. Computer [ÅBE12, AO02, AM10, AJNJ12, AS02a, ABDW10, AF02, AH03, AMGJ12, AJRS03, AN07, AN09, AKG06, AP07, AKR03, AD11, ACR^{+}08, ACD^{+}10, ADR08, AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, BN09a, BBdB12, BHH04b, BHH04c, BPR05, BCE09, BB95, BDW08, BMS04a, BMS08b, BJM09, BI03, BH07, BDZ08, BPZ09, BI11, BCP12, BDSW09, Ber96, BFG^{+}12, BL07a, BC10b, BL03, BPS06, Bod03, BBR88, BBR89, BBR90, BAP^{+}05, BKLS11, BVNW87, BW00, BKM07, BKZ05, BD04, BHK01, BRP09, BEFP08, BtHRBB12, BHH74, Buc85, BLPW02, BN98a, BN98b, BC96b, BC96c, CJ11, Cal82, CGMQ08, CPRS01, CDV03, CRS11, CV04a, CDCW09, Cav85, ČGH^{+}11, Cer07, CP98b, CP98c, CL90, CPC07, CS10, CB06, CSTD08, CJ95, CW88, CKC11, CBBS09, CEER96, Cup00]. Computer [DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, Dat09, Dav89, DFUR11, DRDKE11, DFP^{+}12, DEQ92, DGWB09, DPA07, DH04, DVV07, DR02b, DKN12, DK02a, DK02b, DG05a, DT06b, EDGP11, EKS10, EKF12, Ekl94a, Ekl94b, EP03a, EP03b, EFW93, EMW94, Fau90, yF75, FR01b, FKK04, FMP11a, FMP11b, FMP11c, FJ92, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, Fom06, FM84, FYM12, FTZ08a, FTZ08b, FTZ08c, FTZ08d, FMRW09, FSP09, GP07, GP09a, GP11, GKB^{+}08, GCK06, GM06a, GJRT01, GRS06, Gol94, GKW02, GMS06, GSLM12, GCM08, GW98, GPS03, GKS05, GS10, GCC^{+}02b, GHH06b, GPT10, Gue90, HPTV05, HY12, HS74, HMV01, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, HRSS08, HKLP12, HK10b, HJRS99, Hol07, Hol08, HM09]. Computer [HS11, HNW04, HSL06, HL00, HHG11, HJ00, IN11, JL08, JCC^{+}12, JNPV10a, JNPV10b, JNPV10c, KG12, KP00, hKA10, KKS11c, KKS11d, KKS11e, KKS11f, KSW03, KHG01, KK11d, KW10, KBČ^{+}12, KS10a, KS12c, KU06, KN09b, KNP06, Kuc̆02, KK07, KR11, KV11, KLT09, Kut12a, Kut12b, LNS06a, LNS06b, Lei74, LHH10, LBP06a, LBP06b, LBP06c, LBP06d, LSHB07, LMCH12, LN01b, LYhL^{+}01, LM04, LN04b, LZ02, LRS98, MWD06, MDY^{+}10, MB82, Mah04, MSS11, MLM90, MF08, MP95a, Meh85, MT99a, MLTC11, MAFS08a, MAFS08b, MKR^{+}04, Mit05, MKR07, MVN86, MC89, MDP04, MDBF01, MMK98, MMH07, MS11b, NNS06a, NNS06b, NJ10, NR00, NKM^{+}09, NRR12, NPS06, NSS03, OT08, OS07, OS03a, PM04a, PM04b, PM04c, PM04d, PR03a, PS12a, PH10, PB08a, PMD92, PMD94]. Computer [PMDA96, PMD97, PMDK00, Pin89, PR98a, PR96, Qui98, RRR99, RS06, RFZ03, RT00, RM97b, RV03, RSW12a, SK11, San92, Sch06, SLH04, SMP07, SdDK91, SG03, Squ09, SXC05, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, SKM07, TMLL09, TM09, TY11, Thi10, TW07, TGM97, Tu99, VPBŠ03, VWH06, VBH07, Ver00a, Ver00b, Vet11, WTP^{+}06, YKKZ07a, YKKZ07b, YHR^{+}09a, YHR^{+}09b, YX04, YSR^{+}06, YZCW07, ZiTM10a, ZiTM10b, ZiTM10c, dS02, dS11a, vH83, vLIvdH^{+}07, vLMP^{+}10, vdHHsHD06, vdHS10, vdHP12, vdMvdT08, AI97, ALM95, AN97, AF97, ADL05, AR98, BBEdK98, BB91, Bec75, Bec79, BV92, BS92, BMP90, BC04a, BB05b, BCP99, BC88, BN91b, Bla88, Bli75, BD93, Boh75, BS94, BNS00, BS93, Bra99]. computer [BL01, BLV02, BFJV97, BMS01, BBB98, CV96, Chi95, Cho88, CK84, CJK88, CER79, CLP05, CK82, DK90, DEM81, DDJ00, Dem80, Deu81, DEM96, DPA05, DD05, DN93, DG05b, ENR83, Ehr87, Ehr89, EKR91, EK81a, Eve75, FV01, FV03, FV04, FV05, FHRR05, FZJ01, FMdGC99, Gol82, GT90, GSP01, GS88, GT93, GL03, GP89, GSV05, Gru77, GC81, GRW86, HM94, HP99b, HK92, Heb95, Her92a, HS98c, HM04, Hug91, IM91, Jan88, JS05, JE06, JS84, Jou85a, KSHS01, Kah87, KK87, Ker01, KSS02, KSW05, Knu99b, KP01a, Kra05, KMS89, KU94, KPW99, LBP00, Li96, Mad89, Mah85, Mau91, May93, MST95, Maz76, MT89, MMU96, MR97b, Moh91, Moh97, MP95b, MZF94, MRS97, MO92b, Nag90, Nag95, NDFMS97, NT92, NM94, Nie99, NV01a, Nol81]. computer [Nor86, Nor87, NK88, Nor90, OFK05, OK90, ON80, Pai00, PFCS05, Pat02, PH97, PS96a, Pit86, PPR87, Pit89, Pit91, PRJ95, PZH96, PRR94, PR97, RCA90, RS97a, RVY05, RD84, RRR01, Rol97, RV02, Rov90, SKMR94, SRG04, SE94, SLH05, SPK01, Shy92, Shy93, SC96, Tar91, TC95, TC99, Thi94, Thi95, TS87, TWW77, VBCBS05, WF98, WS02, Wei79, WC98, WNE99, WH95, Win78, YD98, ZBG98, dSG05, ter97, vL89, vL94b, vLWH^{+}00, von79, AM10, ABBC10, AH96, AP04b, ABM02, AFL03, AD11, Aya95, ADK04, BM03a, BN09a, BJ06, BPAB07a, BPAB07b, BS06b, Ber96, BL07a, BKW09b, BTCW10a, BTCW10b, BTCW12, Bor91, Bor92, Bor93, BGM94, BM09a, Bra02, Bra09a, BL02b, Bun96, BRS09, Cal82, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CP98b]. Computer [CP98c, Chi03, CH99, CFVCF^{+}12, CK91, CS00b, CBF01, CCSW12, Cou93a, Cro03, CDT11, CK93, Cup00, DH07a, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, DV10, DEQ92, DPA05, DPA07, DF96, DVV07, Dil94, DH07b, Ehr98, ES00a, Enc80, Ési06, ER05, Fau90, FKP99, FdCLN^{+}11, Fie02, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, FBV11, FR99a, FYM12, FTZ08a, FTZ08b, FTZ08c, FTZ08d, Fri01, FMRW09, FSP09, FMC12a, FMC12b, FMC12c, GP05, GP07, GP09a, GP11, GMV05, GMV06, GMV07, GVT08, GMV09a, GKMV10, GKMV11, GKMV12, Gol94, GKW02, GMS06, GQ11, GPS03, GKS05, Gór06, GGS99, GK09, GHH06b, GPT10, GGK^{+}09a, GGK^{+}09b, Gru97, GM08b, HP99a, HK94, HS08a, HBS00, HLW04, HRSS08, HKLP12]. Computer [HS95c, HV98, HSL06, HS03, Jac07c, Jac07d, Jac07b, Jac07a, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, JL08, JP09a, KS09a, KP06, KM08a, Kap08, KS03c, KKS11c, KKS11d, KKS11e, KKS11f, KH11a, KH11b, KvdM00, KS10a, KS12c, KP15, KMVY06, KMYJ08, KKH07, KV11, LS92, LCHS08, LBP06a, LBP06b, LBP06c, LBP06d, LSY^{+}06, LOS05, LMA09, MS12a, MT04a, MR97a, MJR04, MSWZ06, Mau89, MKZ02, MKZB04, MDP04, MDBF01, MDPQA09, MDPQA12a, MDPQA12b, NNS06a, NNS06b, NT98, NS90, Ong05, OK96, OD12a, OD12b, PT95, PM04a, PM04b, PM04c, PM04d, Pet80, PW03, PMD90, PMD92, PMD94, PMDA96, PMD97, PMDK00, PV04, PGV93, PSS98, PS09c, Qui98, RDPMA^{+}11a, RDPMA^{+}11b, SO07, SRG04, San92, SS01, Sch10b, SLH04, SLH05, SLB05]. Computer [SCL^{+}06, SLL^{+}07, SYY^{+}08, Ska01, SG03, SL99b, SDP97, SKK04, SLF91, TJ08, Tho11, Tom92, TCJ10, VMvN12, Ver00a, Ver00b, Vog01, Wii04, Wii05a, WGD05, Wol95, YKKZ07a, YKKZ07b, YŞ03, YGGÖ05, YKMJ09, ZiTM10a, ZiTM10b, ZiTM10c, ZLJ05, Zip92, dSG05, vB97, vH83, vL95, vP93, BUG93, BGU94, BGU95, CP05]. Computer-Aided [AO02, GW98, HJ00, MDY^{+}10, BHH74, HM04, SC96, CK91, SLF91]. Computer-Assisted [AJNJ12, AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, BHH04b, BHH04c, DK02a, DK02b, DG05a, EP03a, EP03b, FMP11a, FMP11b, FMP11c, JNPV10a, JNPV10b, JNPV10c, LNS06a, LNS06b, LMCH12, MAFS08a, MAFS08b, NJ10, TY11, TGM97, YHR^{+}09a, YHR^{+}09b, DDJ00, DG05b, NV01a, TC99, WC98]. Computer-Human [LCHS08]. Computer-Mediated [PS09c]. Computers [CR04a, CR01b, Dai07, Dwa00, HO03, JS02, Kar09, KF74, Mit06, O’H98, Rob11, TV87, vdV84, HBN06, MF01, MKZK06, MKZK08, MKZK10a, MKZK10b, MKPZ12a, MKPZ12b, SMB03, ZBW94, vdHI99, vdHCD07, vdHXMW08, vdHIP11, Bol78]. Computes [CDL12]. Computing [AP09, AKK04, ASS10, AGP08, AYV^{+}11, Agu12, ACW07, APSZ02, AS11b, APBP07, Ama08b, ASR^{+}07, AFG^{+}11, APP12, ASAM07, AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, ABRS06, BCS09, BHH04b, BHH04c, BCC06, BF08, BP07, BS11e, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BBP^{+}12a, BBP^{+}12b, BWAM09, BDIR07a, BDIR07b, BCC^{+}10, BDG^{+}12b, BMSHW09, BFBS11, BCV06, BBR07, BDDM12, BH02, BKM08a, BEKS02, BUHS08, BS12, BB00b, CYM^{+}11, CPRS01, CCF^{+}08, CLWC08, CDGW10, CB10, CB11, CB12, CL07, CP11, Cer07, CSM10, CGBZ08, CDPT06, CR04a, CH04, CMM09b, CCAS12, CM06, CPPB06, CMCS11, CFP^{+}09, CS12c, DGSM11a, DGSM11b, DL06b, DGT10c, DD03, DSDS04, Dat09, DMS04, DDHP05, DPC^{+}07, DPZ12, DS09b, DsSL^{+}07, DJS06, DSM03, DR09]. Computing [Die03, DSZ10, DMB^{+}07, DLŠ11a, DLŠ11b, DK02a, DK02b, Dol06, DMW06, DF06, DSS^{+}07, DG05a, EKP^{+}07, EP03a, EP03b, EYX^{+}05, FHJ^{+}02, FMZ05, FM04, FL09, Fic03, FMP11a, FMP11b, FMP11c, FSN06, FHY08, FKS10, FPRS06, GMV06, GMV07, GPCL10, GWK09, GY08, GGK06a, GTM10a, GTM10b, GTP12, GRB^{+}07, GMV09a, GKMV10, GKMV11, GKMV12, GK06, GHP^{+}11, GPR^{+}12, GGS09, Gia07, GBC^{+}08, GBC^{+}09, GAA06, GL74a, GL74b, GSS09, GSW06, GHH^{+}06a, GCC^{+}02a, GNTF02, GSV05, Gue04, Han81, HRC06, HG00, HRB12, Her00, HMM12, HTMC09, HPRS06, HYMZ11, HZH05, HLI06a, HLI06b, HLI06c, HHL07b, HHL07a, HNPP07, HKM^{+}07, HWLJ08b, HWLJ08a, HJL^{+}09a, HJL^{+}09b, HZBF10, HZGZ10, HGBF12, HGGG12, HGPH12, HJBF12, HMJG12, HMSU04, ICSY07, IMY^{+}07, IPRO08, Ish10, JZR09, JY96, JM07, JM10]. Computing [JNPV10a, JNPV10b, JNPV10c, JPXS04a, JPXS04b, JGXC04, Jón12a, Jón12b, JLW07b, JL11, JPR04, JDMB08, JDMB10, KEDW07, KN09a, KL95, KPSW10, KMMN11, KWC^{+}12, KLT^{+}12, Kei09, KKM^{+}11, KMS06, KTB09, KHK^{+}08, KVMH09, KS00, KLN07, KSHM09, KASS07, KSG^{+}07, KŚHM11, LLT07, LJS08, LNS06a, LNS06b, LM09, LSDN04a, LSDN04b, LJJG07, LFJI10a, LFJI10b, LJS^{+}10, LCB12, LYCV08, LSK08, LMWY04, LMW08, LMW10, LMW12, LC11, LMC12, LLTY12, LZ05, LTT07, LS10b, LHH11, MJYT06, MRN07, MCK^{+}06, MWYF10, Mah04, Mak86, Mal02, Mal03, Mal07, Mal09, Mal11, MY06, MVMP^{+}04, MN02b, MRY^{+}11, MSAdAS05, MBP^{+}11, MCA^{+}07, MAFS08a, MAFS08b, MCSN10, MPV01, MSB07, MK03b, MBJ^{+}07, MCER10, MSUB04, MSKY10, MNTY05, NO11, ND09, Nie09, NRWI09, NCJK08, OS07, OYM07, OPS11, ORB^{+}09, OWPY03, OZ09, OZ11]. Computing [PPP07, PKKG06, PJC^{+}12, PAD^{+}08a, PDML11, PDSD10, PSDS11a, PSDS11b, PDSN12, PA09, PZYS12, PPJRN^{+}10, Pel07, Pel11, PCS^{+}07, PP03, PISW05, Pri03b, PQ05, PF07b, RCB^{+}04, RMDS10, RR12, RCJ^{+}08, Rau04a, Rau04b, RKM^{+}12, RA00, Rei88, RYG11, RBTD04, RPBP06, RJS^{+}08, RBC^{+}06, RD12b, RSTZ04, RTZŻ06, RTZZ08, RST^{+}10b, RST^{+}10a, RKS^{+}12a, RKS^{+}12b, SPBP08, SW04, SCH^{+}09a, SM11a, SYNC11, SZR^{+}08, Sar03, SMBQ07, SUW02, Sch12, Sch07, SHX^{+}06, SS06b, SAG10b, SMEGA10, SYP^{+}05, SHP^{+}05, Smi06, ST12c, SS09d, SKR^{+}10, Tau08, TPCB12, TG00, THX^{+}07, TBF^{+}09, TSKGB04, UBM02, VJKT00, VJAA03, WWD99a, WWD99b, WBHR08, WHR10, WCBD08, WYX08, XYM^{+}07, XBS^{+}10, YGGJ04, YJMU06, YHR^{+}09a, YHR^{+}09b, YMHH10, YGE12, YYS^{+}12, YKM06, YLC^{+}10, ZPTI09, ZCDT10, ZJL06, ZSY^{+}06, ZZLL10, ZF05, ZJS^{+}02, dLWC03, ABS01]. computing [AM91a, AFK90, AAN00, BPSP06, Ban92, Ban93, Ban94, BKS99, BBKP06, BL05b, BIRT05, BB95, BCT06, BB05b, Bod91, Boe01, BU01, BTF02, BMYZ95, BP04b, Bre97, BAWH00, BLS02, Cag00, Cag02, CC93, CP06a, CS11, CF00a, Cha05a, CPC^{+}99, CL04, CR01b, CS97, CFP94, DKP05, DB02, DS05b, DFK91, DDJ00, DMWW88, Dik04, DGR92, DW94, DMW96, DG05b, EHP94, ELM05, FL00, Fog94, Fog95, Fog96, Fra05b, GMV05, Gel99, GWS05, GH94a, GH94b, GM04a, GZR02, HO03, HI93, HK98a, HS00, HPR05, HS95b, HS97, HHW01, HMP99, HSS96, HGZ05, HK96, HVT12, HO98, HSB^{+}96, HW05, HU05, Ish99a, Ish97, Jay99, JM02, JRJ05, JS02, Kåg98, KKKN05, KS05a, Kut98, Lee01, Lem97, Li98, LCHS96, LSS^{+}04, LMW06, LA05, Mal95, Mal97b]. computing [Mal99, Mal01, Mal05, MAD^{+}04, MWY01, MMT97, MKT98, MPPJ^{+}05, MMG^{+}01, M^{+}06, MSS05a, MBA05, NOOS05, NV01a, NA97, PS02a, PDHS03, PS97a, PS98a, Par02, Pau03, Pin95, PS98b, Pol97, Pol99b, PT06b, Rai03, RBC^{+}05, SKS02, Sch93a, Sch03a, Seh97, SU97, SL99a, SBH98, Slo99, SMMG01, TTP05, TC99, TWN92, VKV00, Wan03, Waś96, Wel01, WC98, Wil99, X^{+}12b, YAL^{+}05, YRMD05, YT99, ZSO99, ZY01, ZVU99, vZvdV89, Asa99, yCW96, Cai11, COT98, CL06b, CM04a, DL95b, Du00, FD11, GTM09a, GMV12, HK98b, HW08a, IZ02, JL97, KKW11, KKW12, Lin07, MTO99, Ngo09, O’D77, SdDK91, SWS^{+}05, TS10, Wan01, Wan05b, WZ03, WAW01, WH11, Pad99, San80a, SUW02, SS90]. computing-ARCS [SUW02]. Comwim [WTZ^{+}04]. Concept [Ber12, Dau03, DIP12, Ekl04, FR09, GSW05, KS07c, KS10c, MO08, MS06b, VJ11, GG05b, Ekl04, YNB08]. Conceptions [CR05b]. Concepts [BTPPS06, BTPPS07, BTBP^{+}08, BTPPS09, BTBP^{+}10a, BTBP^{+}10b, BTKP^{+}11, BTPP^{+}12, Bod03, BW00, BKM07, BEFP08, DBSvD09, Dig11, Fom06, GSLM12, HRR^{+}06, HNW04, HKV10, Jae01, KK11d, KKMN07, Kuc̆02, Mel04, MYS10, PH10, Som09, SPS09, Thi10, THR03, WL04, BS92, BTPPS05, BL01, DR81, GK01, GS88, HS98c, JHC02, JM78, Kra05, May93, MST95, Moh91, Moh97, Nag90, Nag95, NDFMS97, Nol81, NA97, Sch97a, TS87, Van94, WNE99, vL89, vL94b, Bee86, Bee89]. Conceptual [ABCWM99, APHA11, CVLL12, CW07, CFL10, CT99, DMB^{+}11, DKM^{+}05, DS01a, EEM96, EH08, Ell95, EG97, EOR06, GM00, HRK^{+}07, HP09, KD11, KJS01, LLS00, LCD^{+}09, LSYO08, LMT00, LRL98, Luk97, MMS93, MC98, OYY03, PSST07, PSS^{+}10a, PCA02, PPH07, RBsC^{+}06, RDK09, SHØ06, SLLS03, SPC^{+}08, SMK02, SDM05, TDS94, Tha96, TDK^{+}10, WPD04, dMLG03, ALS^{+}05, ACL^{+}04, BJM^{+}05, Che99a, ABCWM99, APHA11, Ari02, ACL^{+}04, ACR12, BCGY09, Che99b, CFL10, CT99, DKM^{+}05, DS01a, DKS12, EEM96, EH08, Ell95, EG97, EOR06, GM00, JP03, JDL11, KJS01, LLS00, LCD^{+}09, LSYO08, LMT00, LRL98, Luk97, MMS93, MC98, PSST07, PSS^{+}10a, PN93, PCA02, PPH07, RDK09, SHØ06, SLLS03, SMK02, SDM05, TDS94, Tha96, WTZ^{+}04, WPD04, dMLG03]. concerns [YM01]. Conclog [Jac91]. Concrete [Kle89]. CONCUR [AD05, BH06a, BZ09, CV07, GL10a, KK11a, KU12, vBC08, AD05, AL03b, BK90, BG91a, BH06a, Bes93, BZ09, BJKK02, CV07, Cle92, GY04, GL10a, JP94, KK11a, KU12, LN01a, LS95, MW97, MS96c, Pal00a, vBC08]. CONCUR’98 [Sd98]. CONCUR’99 [BM99a]. Concurrency [AD05, AL03b, BG91a, BH06a, Bes93, BZ09, BJKK02, BRW85, CV07, Cle92, DRR04, DT06a, GY04, GL10a, JvdAB08, JvdA09, JBK09, JDK10, JDK12, JvdAM^{+}12, KK11a, KY07b, KU12, Vog88, vBC08, BK90, BM99a, BRR87a, Cas81, JP94, KL95, LN01a, Lau93a, LS95, MW97, MB96, MS96c, Sd98, Zwi89, dBRdR86, dBdRR89, dBdRR94, CYR02, DHS10, DDM08, JY96, Pal00a, Vog88, YI91]. Concurrent [BR97, CNY95, Den85, God96, Gue90, Hai82, Jac91, Kah79, TWN92, ADR01, Bee00, Har77]. Conditional [DL95a, KJ88, KO91, RR93]. Conditionals [KI01, KI01, KIRK05]. conducted [BZ83]. Confederated [MTP^{+}02, MT05a, MTH05, MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MTH08, MDH09a, MDH09b, MHD09, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, MTS^{+}03, MT^{+}03]. Conferencce [SDP97]. Conference [ÅBE12, AH00, AO02, AAW04, AG01, ACdR^{+}08, AD05, APFV09, AB10, ARS03, ARS09, AK07, AP09, AL11a, AGB^{+}11, AKZ01, Abe06, Abe10, ABB12, APR^{+}06, ACN^{+}07, AK90, AV99, AJNJ12, AMS06, ABMM07, Abr07, ALS^{+}11, AM91b, AM91a, Abr01, AS11a, AI06, AB11, ASS10, AJRB09, ABH12, AST05, AK97, AAA^{+}11, AT02a, ABK^{+}09, AFP^{+}10, AS02a, ADDL08, ACL12, AW12, AYV^{+}11, ABG09a, AW01, AB12, ANS04a, ANS04b, ANS04c, AVP07, AKU01, AK03, AKT06, ACK^{+}07, ABKT11, ACD^{+}07, Akm01, ABCWM99, AM97, AMU03, ASE^{+}07, ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12, ASP01, AKM10a, AKM10b, ADJ^{+}01a, ADJ^{+}01b, AvAS06, AvASD06a, AvASD06b, AvASD06c, AL04, AGL^{+}07, AE05, AD06, APPE09a, APPE09b, AV07, ABB^{+}10a, AF11, All96, ARC^{+}07, AH02, Alo05, ADR07, ABB10b]. Conference [APSZ02, ABB99, ABMH08, AJB09, AS11b, AL03a, AP04b, APBP07, AL03b, Ama08a, ACR07, AC11, ASLM09, Ame91a, ABD^{+}99, ASR^{+}07, ALPH10, AFG^{+}11, AFS02, AHK07, ABS91, AFGvE10, AFGvE11, ACH00, ABM02, AFL03, APW06, ABJ02, ABH^{+}10, ACL98, AMCL02, AYB^{+}09, ABK12, ATO^{+}09, APHA11, ALR11, AIL12, AS98, APS06, AGS^{+}11b, AGS^{+}11a, Ant97, AP11, AT12, AJ11, ABG96, AHJ07a, AHJ07b, AHJ07c, AGI^{+}06, AMPT09, AT02c, AS10, AS12, ABGL09, ABM05, AM98, AF99b, AM00, AJS00, ABD08, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, AWA^{+}11a, AWA^{+}11b, AKG06, AR98, AP07, AHMT08, Asa99, AB06, AB03, ABD03, ABT04, Asp98, ASAM07, ALZ11, Atl08, AJ01b, ACJ08, ACR12, APM08, AM05, ADP11, AJN12, Aus86, ACR^{+}08, ACD^{+}10, AKTT12, AKSO06, AOM07a, AOM07b, ADS09, ADGM12a, ADGM12b, ADGM12c]. Conference [Ayk07a, Ayk07b, Ayk09, ALL^{+}10, AKR^{+}09, BBE01, Baa03, Baa07, BSMM04, BV02, BM03a, BVO10, BG10a, BW05a, BE11, Bac00, BO00a, BC09a, BBCL08, BG91a, BM99a, BBdB12, BYdVZ^{+}12, BS11a, BA06, BW03a, BH06a, Bai05, BMS^{+}08a, BBT08, BHL^{+}03, BMK08, BMK09, BDK10a, BKH11, BC95, BLS06, BBGS10a, BBGS10b, BBGS10c, BST01, BST03, BJ06, BTT90, BB07, BCT02a, BMUV10, BKS99, BS11c, BR12a, BCD^{+}03, BLO^{+}07, BLW09, BYLJ11, BW11, BSZ12, BPP91, BSW95a, BA99, BGLT05, BBS07, BPAB07a, BPAB07b, BKA11, BH06b, BFH07, BCS09, BHH04b, BHH04c, BPCS06, BA07, BKLM07, BMPM09, BHKR11, BG93a, BFF^{+}10, BGK12, Bar03, BB90, BdB08, BH10, BPS11, BL06a, BW03b, BR04, Bat01, BCT02b, BS02a, BSN06, BCL03, BS11d, BS11e, Bat09, BW10a, BGV01, BMP00, BMP01, BTQ02]. Conference [BTQ03, BBP04, BPW06, BCE09, BD12, BHHK08, BS06a, BPV04, BPR08, BH08, BM11a, BDG12a, BBLT06, BDL08, BB99, BL05b, BCRB^{+}11, BB95, BJP12, BDIR07a, BDIR07b, BH06c, BPR02, BL08, Bel00, BP11, BCC^{+}10, BAHH07, BR92, BGKY08, BEP04, BD97a, BDCM04, BCG^{+}07, BCKR10, BCST99, BDG04, BB01a, BG11a, Ben06, BV92, BLJ08, BCD05, BDd09, BMSHW09, BFBS11, BdR94, BF09, Ber10, BMS08b, BJM09, BNUW09, Ber08, BLKP08, BF12, BFGL10, BKH^{+}09, BL10a, BC11a, BF11, BC06, Ber98, Ber02b, BBKW03, BLB^{+}03, Ber00, BCP^{+}04, BDH^{+}99, BCP99, Bes07, Bes93, BKL^{+}12a, BKL^{+}12b, BB01b, BND01, BG93b, BHCL87, BKW08, BKW09a, BW10b, BK80, BM10, BD97b, BFG^{+}12, BN04, BG06, BCB97, BS01a, BG03b, Bih03, BR96a, BZU07, Bir12, Bis02]. Conference [BV11, BH92, BBKM11, BF00, BFP04, BBRM98, BBP96, BBZ00, BBZ01, BSS11, BV12, BSBM05, Bla03a, BMS04b, BTPPS06, BTPPS07, BTBP^{+}08, BTPPS09, BTBP^{+}10a, BTBP^{+}10b, BTKP^{+}11, BTPP^{+}12, Bla03b, BS02b, BRST08, BB10, BL03, BPS06, Bod91, BRW93, BLKW01, BR12b, Bod05, BGPW05, BI73, BKLP11, BKLR12, BFSN10, BCC^{+}12, BM02a, BDS04, BKW09b, BTCW10a, BTCW10b, BMW10, BTCW12, BG10b, BDK10b, BTZ^{+}10, BO00b, BKS01, BI04, BKS05, BLvdH10, BTF02, Bon03, BGP00, BDDM12, BJ07, Boo91, BB00a, BNS00, BBBB08, BCC07, BH02, BF06, Bos01, BK04a, BL10b, BHJ^{+}11, Bos96, BS03b, BF10, BM09a, Bou11, BMYZ91, BMYZ95, BMCCM11, BCRT92, BFMR96a, BFMR96b, BBMSD09, BKM08a, BHL11, BCK10, BEGP04, BEG04, BJ01b, BIMW12, BSB^{+}99, BZ08a, BAP^{+}05, BKM01, BGKS04]. Conference [BH95, BHT97, BFH98, BDGK00, Bow94, BDGN07, BCD00, Boy01, BM03c, BN09b, BC11b, BKLS11, BH05, BR07b, BR09, BDV95, Bra02, BRT11, BTT12, BLV02, BvdHPP09, Bra73, BCF06, BZ09, Bra93, BJ01c, BW01, BC09b, BGHS05, BKW06, BHN97, Bri93, BJKK02, BHQ10a, BHQ10b, BUHS08, BL02b, BRP09, BP12b, BZ03, BJ12b, BtHRBB12, Bru06b, BD11, BM11b, BHH74, BSS00, BCR03, BG07b, B^{+}12b, BAWH00, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BvADS08a, BvADS08b, BMC08, Buc85, BV94, BBL87, Bud91, BHH08, BOH^{+}12, Bun94, Bun96, BZ08b, BGMST09a, BGMST09b, BCC99, BC98, BE01, BD03, BR07c, BLVV07, BN98a, BN98b, Bur90, BTMI11, BL06d, BLS02, BRS09, BPS02, BHLP07, BL11, BC96b, BC96c, CSPC09, CRJ11a, CRJ11b, CV11, CC04, CCL06, CV07, CSCO00, COBV11, CFI06]. Conference [CD10, CG07, CYM^{+}11, Cal82, CC93, CC95, CP98a, CBC^{+}02, CIW06, CDP02, CDV03, CCD04, CDP^{+}06, CCF^{+}08, CCD^{+}09, CHM^{+}10, CKPR11, CLM03, CL08a, CKS04, CRL01, CK04a, CK04b, CK06b, CK06a, CK08, CK10a, CK10b, CK12b, CK12c, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, Can08, CV04b, CPFD^{+}03a, CPFD^{+}03b, CSJD06, CLWC08, CFZ10, CZF10, CE01, Car03, CS11, CDCW09, CP10a, CNP^{+}08, COMTLB12, CMT02, CVTP12, CJ08a, CFGN11, CIT11, CFH^{+}12, CD07, Cav85, CL07, Cer05, ČGH^{+}11, CC07, CD00a, CGP02, CCPP12, Cer07, CVV08, CBNR02, CCJ07, CT07, CSM10, Cha05a, CCD^{+}06a, CCD^{+}06b, CM03a, CG02, CGBZ08, CCL02, CKK^{+}09, CHCM11, Cha02b, CH10, CDPT06, CMM^{+}09a, CW09, CLY02, CCH02, Che03b, CCM^{+}04, CH04, CL06b, CDLZ06, CJ08b, CRW08, CMS08a, CC09a, CMM09b, Che10]. Conference [CY10, CTYG10, CTS10, Che11, CY11a, CC12, CYZ12, CP06b, CLZ06, CKLN10, CEQ^{+}11, CHG^{+}07, CNA08, CWL01, CCdS03, CFK^{+}12, CHCA09, CY12, Chi95, CP98b, CP98c, CJ05, CJ06, CH09, CPC07, Cho02a, Cho02b, CK06c, Cho12, CRD08, CKH10, CCBW11, CDCJL08, CH99, CS05b, CHAL04, CDY^{+}11, Chr91, Chr00b, CKM^{+}09, CHA03b, Chu03, CM06, CM04a, CH96, CW99, CW06, CGP03, CS12a, CP10c, CFVV^{+}12, CLB11, CPPB06, CKM93, CCK95, CA10, CV10, CVV12, CU08, Cle92, Cle99, CS00b, CFG^{+}11, Cod01, Coe98, Coe11, CGK^{+}12, CCD^{+}12a, CCD^{+}12b, Coe02, CSTD08, CS02a, Coh10, Coi96, Coi99, CH91, Col02, CTE^{+}06, CML^{+}10, CK01, CSAH09, Com97, CBF01, CMP07, CDSS07, CUdlC04, CMV96, CS01a, CNmWH^{+}05, CR73a, CR73b, CGGO03, CS02b, CP07, CLS07b, CK07]. Conference [CDL12, Cop95, CYBF06, CWO^{+}09, CY09, CRS10, CKW11a, CKW11b, CSA^{+}12a, CSA^{+}12b, CBW99, CEKR02, CEM^{+}06, CKC11, CTS07, CR04b, CCSW12, CC97b, CvH07, CC09b, CSRS08, Cou05, CMB^{+}08, CM10, CPR12, CF10, CS01b, Cra05, Cra12, CP02, CBBS09, CPP07, CDS09, CR05b, CFS06, CGB11, CFL10, CFRR10, Cro03, CDT11, CDA^{+}06, CK93, CDG89, CGBDA08, COB06, Cue03, CMQ96, CHR10, CBCGA12, Cur09, CFB11, CF00b, CD11, CD12, CT99, COB^{+}08, DGT10a, DGT10b, D’H10b, DGSM11a, DGSM11b, DFL^{+}10, DGM08a, DGM08b, DN07, DSZ04, Dai07, DH05, DL06a, DYCD09, DH07a, DL06b, DLV04, Dan12, DF10, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, DS05a, DVVA08, DG04, DPLW08, Dat09, DSZ^{+}11, Dav89, DFUR11, DF01b, DFR02, DSM94, DSB04]. Conference [DMS04, DJAP07, DG07a, DV10, DB98a, DW07a, DEKR09, DDHP05, DPC^{+}07, DKS^{+}07, DKC10, DSB10, DLNU11, DC12, DH97, DMR01, De 07a, DR11, DD95, DFM04, DPZ12, DY06a, De 10a, DG02, DZ10a, DS09b, DPB^{+}04a, DPB^{+}04b, DBC^{+}10, DRDKE11, Dec00, DGG^{+}11a, DIMV09, DFMV10, DMV12, DG09a, DG12, DFK91, Dei90, DKM^{+}05, Del00, DF11, DKM91, DDJ00, DS01a, DQBZ02, DBPZ05, DWWL09, DMLW11a, DMLW11b, DMLW11c, DBB^{+}08a, DV07a, DFG^{+}12, DGLT12, Der89, DV07b, DLW12, DS98, DPW04, DH10a, Des09, DSM03, DKLT05, DCP95, DR06, DY10, DB09, DDI10a, DDI10b, DDI10c, DH94, DO89a, DO89b, DPA07, DD10, DN09, DC00, DJJ12, DD07, DTMH10, Dik04, DS08, DBHS10, Dim03, DDK11, DSZ10, DG09b, DKU05, DGN01, DFN97, DW10, DLŠ11a, DLŠ11b, DPT08, DK02a]. Conference [DK02b, DKB03, DS11c, DIP12, DFF06, DFPS06, DFS08, DFM11, DFT12, DPMP^{+}11a, DPMP^{+}11b, DA08, DT06a, DGG^{+}10, DTW03, DW03b, DLW^{+}07, DZ10b, DP97, DPNM12, DBH01, DBB^{+}08b, DBD^{+}10b, DBB^{+}12b, Dor02, DF06, DW07b, DXZ07, Dro09, Du00, DHP09, DFW^{+}11, DP06, Dub09, DKOZ05b, DKOZ05a, Due04, Duf07, Duf09, Duf11, DHH^{+}08, DRS08, DG97, Dun12, DH07b, DLM07, DLJ12, DFS06, DKW07, Dwo06, DL07, Dyc00, ENP02, EV96, EJM10, EOE^{+}07, EHP94, EB04, ERW99, EM03, EHKP05, EBT11, ELS12, EK09, EM02, EGMW11, EF02, Ehr98, EFG^{+}09, Ehr85a, Ehr85b, EEK^{+}87, EKLM87, EHRT08, ERRS10, EEKR12, EBSS98, ES10, EFT01, EEM96, Ekl04, EH08, Ekl94a, Ekl94b, ECB06, EHH12, EM06, EK10, Ell95, EP03a, EP03b, EKT94, ELNM08, ELN^{+}10, EML^{+}12]. Conference [EMT02a, EMT02b, EHM11, EG97, EOR06, Emb98, ES00a, EN06, EW04, ES03, EOSW07, EBT07, EKO^{+}04, ELS09, EW03, Ern07, EP07, ETG02, EF03, Ési06, EAES^{+}10, EL02b, EM10, EBAH08, EV09, EVV^{+}11, Est98, ET06, ES92, ES00b, EKM06, ED06b, EKS00, ER12, FRvKS10, FHJ^{+}02, FT11, FSA^{+}07, FWY05, FGH98, FK10, Fau90, FZJE^{+}05, FZJE^{+}06, FMC04, FM97, FS07b, Feh11, FR11, FKK09, FKP99, yF75, FLY06, FSM03, FV10, FGB11, FSdlPT11a, Fer08, FR09, FLMG10, FM04, FI08, FdCLJ^{+}11, Fic03, FMP11a, FMP11b, FMP11c, Fie02, FV09, FOGE^{+}09, FD93, FPCK11, Fin99, FNY93, FHFL06, FHLP09, FHKQ12, FS99b, FTAZ06, FW07, Fis09, FBM12, FvdHKL06, FSN06, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, FDC12a, FDC12b]. Conference [FBV11, FK12a, FB09, FX08, Fle02, FLF^{+}08, FMN99, FKS10, FR99a, FSV09, FCLL02, FKW08, Fon00, Fon03, FRL07, FG96, Foo99, FP08, FJ10, FGK^{+}11, FTZ08a, FTZ08b, FTZ08c, FTZ08d, Fos02, FNPW05, Fra00, FKBP11, FKBA12, FW09, FKM07, FC93, Fra01, Fra99, Fra04, FHW08, FGK^{+}10, FL11, FG10, FGM92, FGL96, FSTS09, FT04b, FLL03, FCLG04, Fre98, FM99, FKS07, FNGW08, Fre96, FGTZ08, FLR11, Fri01, FSP09, Fri94, FK95b, FD11, Fum97, Fur05, FS06, FN12b, FKL08, FLP11, FSS06a, FSS06b, FTF91, FKLY08, FTC^{+}10, GO96, Gab97, GHJ06a, GHJ06b, GHJ06c, Gac02, GGD09, GP07, GP09a, GP11, GPAS^{+}06, Gal07, GP08a, GN12a, GTT04, Gal97b, GCB^{+}05, GCJB09, GLP12, GE08, GM00, Gan99, GVM99, GML^{+}08a, GJ09, GLSY10, GLW^{+}12, GH03, GLMP07]. Conference [GMO09, GD10, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GY04, GWK09, GRT02, GS01, GLV09, GVT08, GL10a, GJ93, Gau93, GFV00, GJT03, GLZZ09, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, GGK06a, GHR89, GLCT10, GKC^{+}12, Gec81, GF03, GKB^{+}08, GPAD08, GT03, Gel01, Gel02, Gel03, Gel04, GdATM05, Gel06, GRG06, Gel07, GM07a, Gel08, GM08a, Gel09, Gel10, Gel11b, Gel11a, Gel12a, Gel12b, GLP99, Gen09, GH94a, GH94b, GRB^{+}07, GM03, GK06, GGHN97, GL10b, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, GG07a, GG07b, GG07c, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, GHP^{+}11, GPR^{+}12, GM89, GGS09, GDP07, GVB12, GO11, GAA06, GO09, GN12c, GW09, GR12, Gie05, GH10, GMBM05, Gil10, Gir99, Giu98, GT04, GN88, GZR02, GK12, GH09, GSH^{+}04, GMI06, GMUW12, GK95]. Conference [GP09b, GJN10, Goe98, GK01, GW11a, GCSR07, GZ09, GLIC10, GF12a, GPB02, GPE05, GPYD09, GHSY06, GG10, GLC^{+}11a, GLC^{+}11b, GdlP99, GTVC06, GW98, GQ11, Gor03, GLN01, GS95, GW06a, Gór06, GH96, GH12b, GR06a, GV95, GGS99, GS10, GK09, GBMS11, GS00b, GMS12, GSW06, GYA^{+}06, GL92, GJS08, GHH^{+}06a, GL08, GCC^{+}02b, GTY03, GSS00, GL03, GS08, GP02, GGK^{+}09a, GGK^{+}09b, GAM^{+}11, GW12, GH04, GSV05, GH07b, GMV09b, GKMN10, GW93, GN98, GDGB12, GV09, GMV12, Gue99, Gue01, Gue04, GQ95, GW11b, GHMV11a, GHMV11b, GHMV11c, GF97, GKN03, GM08b, Gup10, GM10, GM07b, GHK^{+}10, GW00, Gyi96, GPV88, HM11a, HP12, Hae99, Hal94, HP99a, HR06, Hal09, HS12b, HT97, HCT02, Ham08, HT09, HMT10a, HLSZ11a, HLSZ11b]. Conference [HT11, HHMT12, Ham00, HS07, HTW02, HRdV11, HMS09, Han81, HHJ^{+}86, HWP01, HLR^{+}98, HLC^{+}12, HM12a, HB97, HAM95, HMV01, HKP08, HKL07, HHLS08, HRC06, Har07, Har09, Har11a, HS08a, HMTH10, HG00, HK05a, HZ10, HKI^{+}09, HS08b, HBS00, HLMW06, HM12b, HS00, HLKX12, Hed03, HNM02, HKP10, HU08, HLW04, HMW^{+}08, HJ12, HW01, HJMP06, Hen08, HA02, HWG10, HK10a, HBN06, HRB12, Her00, HV06, HP06, HC11, HMM12, HSP^{+}12, HAB98, HHJZ12, HBE07, HS95b, HHW01, HG98, Heu98, Heu00, HN12, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, HK11, Hig05, HW09, Hir97, Hir98, HF97, HL10, HS95c, HJW00, HMP99, HSS96, HCL05, HK05c, HZ08, HGZ05, HW07, Hof97, Hof01, Hof03b, Hof11, HCR06, HW03, HK96, HLW08, HSB^{+}10, HŽ07, HB11, HS11]. Conference [Hon01, HDGK11a, HDGK11b, HM01, HM05a, HRM^{+}10, HMMW10, HVT12, HRRY08, HNP12, HSL^{+}99, HSH08, HCDL09, HH02, Hor02, HY03, HNS^{+}12, HMPZ09, HPP88, HKV10, HO98, HK98b, HYPY10a, HYPY10b, HYMZ11, HV98, HW08a, HLCZ11, HSSV12, HM12c, HdL12, HC09, HZH05, HLI06a, HLI06b, HLI06c, HHL07b, HHL07a, HWLJ08b, HWLJ08a, HXC^{+}08, HYP^{+}09, HJL^{+}09a, HJL^{+}09b, HZBF10, HZGZ10, HCS11a, HCS11b, HGBF12, HGGG12, HGPH12, HJBF12, HMJG12, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, HSMPA09, HP89, HL00, cHPkC^{+}07, HS99, HMT10c, HKH10, HLFS12, HJ00, HS03, HP99c, Hus01, HS98d, HDLM09, HMSU04, HU05, HST07, HSVZ10, HKM^{+}11, IdKM06, IZ02, ID03, ID06, IY06, IR08, Ibr95, IKR00, ILR95, IAC06, IRM93, Ima99, IR07, IMY^{+}07, IPRO08, IL01, IJ06, IFGH10, IFL^{+}12, IKY05, ISS^{+}06, INR07, IAL^{+}07, IK12]. Conference [Ish11, Ish97, IDMY08a, IDMY08b, IS02, IS12a, IS12b, IS08, IM91, IT03, IKKU08, IT08, IZP09, JZR09, Jac07c, Jac07d, Jac07b, Jac07a, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, JNS05a, Jac04, JY96, Jaf99, Jäh99, JSV11, JW05, JM05, JM11a, JYZ04, JLY03, JN10, JRL09, JM07, JM10, JSWhT11, JS01, JT09, JBJ94, JO99, JKL02, JS97, JFBH08, JM02, JNR11a, JNR11b, JZ03, JS08, JNH11a, JNH11b, JJP^{+}02, JP04a, JPD04, JLS11a, JLS11b, Jeu98, JCC^{+}12, JDL11, JHC02, JKN^{+}12, JM11b, JS11a, JWH06, JP09a, JNPV10a, JNPV10b, JNPV10c, JDAW12, JWbG^{+}06, JWG^{+}06, JPXS04a, JGXC04, JRPP07, JM03, JV06, JP11a, Jón12a, Jón12b, JMO09, JTB^{+}05, JMK09, JC12, JMP12, JS11b, Jou09, JY14, Joy03, JMO10, JMT11, Jue04, Jul98, JK06, JN12]. Conference [KSG^{+}09, KG12, Kah03, KG04, KPS12, KM11, KMTW03, KP06, KPK05, KMT00, KWA01, KWA02, KMW03, KMW04, KC07, KC09, KKGK11, KC11a, KC11b, KM08a, KKS11a, KKS11b, KJR05, KKKS11, KL95, KLZ07b, KA12, KUA00, KPSW10, KMR^{+}05, KL07a, KSS08, KP00, KMMN11, KvETvdH10b, KvETvdH10a, Kap92, KR02a, KR04a, KJ09, Kar77, Kar83, KV04, Kar09, KL07b, KP92, KS02c, KK11a, KLP04, KLB^{+}06, KLS10, Kat92, KY07a, KWC^{+}12, KKMW07, KP10a, KLT^{+}12, Kay03, KP04, KLM^{+}07, KOL^{+}04, KS01, KS03c, Ker97, Ker01, KBP07, Ker12, KMS^{+}11, KLP10, KS12b, Kia11, KL12, Kil01, KM96, Kim01b, Kim02, Kim05b, KS06, hKhLKŚ10, hKAS^{+}11, Kim12, KKS11c, KKS11d, KKS11e, KKS11f, KM04, KWCW06a, KWCW06b, KWCW06c, KjHGG11, Kir93, KK98, KR02b, KILW10a, KILW10b, KN03, KS02d]. Conference [KSUZ11, KD99, KY07b, KS93, KGG^{+}11, KO11, Kno11, Knu01, Knu02, KN83, KLLP12, KW07b, KS04b, KFW04, KKdM^{+}07a, KKdM^{+}07b, KRRBV07, KTB09, KHL10, KSDO06b, KSDO06a, KBHSW06, KHSM^{+}08, KK11e, KHW03, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KCMO11, KLPP03, KPK^{+}10, KC02, KvdM00, KFMM06, KKC09a, KKC09b, KRH09, KV08, KK09b, KFLS07, KBH03, Kos98, KI12, KKMN07, KS10a, KS12c, KHI06, KHS07, KMMT12, KU12, KFM07, KP09, Koz04, KP05, KLN07, KO07a, KKL12, KT11, KLFK10, Kra98, KJ10, KSHM09, KO07b, KP10b, KPR07, KP11, KP15, KOS07, KKH07, KS12d, KF74, KASS07, KNK08a, KNK08b, KS91, KPRS07, KSTP99, KNP06, KMN94, KWT03, KSGT10, KRS02, Kul09, KKGT03c, KKGT03a, KKGT03b, KZ10a, KR12, KMMP12, KJS01, KR06, KSG^{+}07, KMS10, Kur07, Kur10]. Conference [Kur09, Kur11, KLT09, KS07c, KMKP11, KŚHM11, KDWW10, KFfKL11, KS10c, LLT07, LLS00, LCD^{+}09, LGK^{+}04c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, LC06, LYL11, LZL11, Lai03, LMR^{+}06, LJR^{+}10, LOX99, LPT07, LK10, LM98a, LM06a, LLW06, LSS02, LV07a, LLMT04, LN01a, LPOA^{+}06, LNS06a, LNS06b, LLCL05, LB05, Lau12, LK99, Lau00, LW95, LGTB03, LGBT03, LZ08, LP97, LOR10, LLP99, LS95, LM98b, LT00, LL03, LhK03, Lee04, LLWL04, LTW06, LL07, LKK^{+}07, LCHS08, hLhKcFŚ09, LLPH09, LC09, LBM^{+}10, LH10, LW11, LTL^{+}11a, LTL^{+}11b, LHŚ11, Lee11, LHKŚ12, gLPZ^{+}12a, gLPZ^{+}12b, LWDM12, Lei74, LM06b, LRG04, LGG97, LBP06a, LBP06b, LBP06c, LBP06d, Leo00, LG99, LVP04, LwCM00, LLLL08, LLC09a, LLC09b, LPC^{+}11, LW09, LSE02, Li96, LLT^{+}04, LWF04, LZhL06, LFIM07, LLIH07, LJJG07, LZS^{+}08, LSYO08, LKZ^{+}08, LHDD08, LMA09, LFJI10a, LFJI10b, LJS^{+}10]. Conference [Li11a, Li11b, LCB12, LPR12, LNW^{+}12, LCF^{+}04, LSTZ12a, LSTZ12b, LN03a, LS06a, LP10a, LFK^{+}02, LL04, Lin07, LTW11, LXH^{+}12, Lin12, Lin14, LRL98, LJ00, LPH00, LNS11, LM03, LYL06, LSK08, LMWY04, LMW08, LMW10, LMW12, LN01b, LYhL^{+}01, Liu01, LCY03, LLL06, LH06, LMA08, LWYN09, LBDW09, LLWL09, LDXZ10a, LDXZ10b, LC11, Liu11, LMC12, LLTY12, LS04b, Llo00, LM04, LPR08, LMT10, LRH10, LPA00, LPZ08, Lom93, LST^{+}12, LWRdR11, LLMR09, LPK05, LT11b, LP00, Lor01a, Lor01b, Lou92, Lou94, Lov82, LHJ08a, LHJ08b, LHJ08c, LNSS11, LGM11, LZ00, LZ05, LMM10, LZK11a, LZK11b, LZK11c, LM08b, LS06c, Luk97, LTT07, LKK10, Luo04, Luo06, Luo07, Luo08, Luo09, LSW^{+}10, Luo10, Luo11, Luo12, LMB08, LO88, LHH11, MJYT06, MRN07, MOP11, MOH^{+}12, Mac12, MOP^{+}08, MRC12, MT09a, MS12a, MCK^{+}06, Mad92, MWYF10]. Conference [MPV12, Mag02, MSB12, MP98, Mah04, MPP08, Mai00, MSG07, MG08, MF11a, MF11b, MMV06, MT09c, Mal02, Mal08, MBS11, Mal99, Mal01, Mal03, Mal07, Mal09, Mal11, MPF95, ML04, MAD^{+}04, Man09, MBW08, MN02a, MEK03, MPS06, MS03a, MSS11, MMR07, MPdSM07, MM08, MT04a, Mar11a, Mar11b, MM01, MY01, MWY01, MVW09, MR01, Mar05a, MRŠ03, MMP03, MMV04, MPB05, MVC07, MSZ09, MVL11, MOBS06, MdRIR08, MSS07, MdlBP05a, MdlBP05b, MW11, Mar93, MNW01, MA03a, MF01, MBMS07a, MBMS07b, MLM90, MVOF08, MSW07, MTCOK10, MTCOBH11, MM89, MBVTY11, MMDN12, MKT98, Mat01, MM03, MM07, MM09b, Mat09, MSWZ06, MN02b, MS03b, Mau89, Mau96, MW03, ML11, MP03b, May01, MW97, McA00, McC96, McC97, MBP^{+}11, MG07a]. Conference [MT04b, MW09, MPvH04, McK99, MS02a, MS96b, MO08, MZC06, MMO^{+}11a, MMO^{+}11b, Mei08, MT08c, Mel07, MSS03a, MS02b, Men05, Men07, MSW02, MN98, MAB^{+}02, MHA09, MM10, MH11, MR08, MMG10, MAFS08a, MAFS08b, MA11, MJ07, MNW08, MW08, MM04, Mez11, Mic10, MSH08, Mid01, MKZ02, MKZB04, MKZK06, MKZK08, MKZK10a, MKZK10b, MKPZ12a, MKPZ12b, MHK05, Mil12, Mil00, Mil01, MP93, MCSY07, MMS93, MMMD97, MdPA98, MSA99a, MSA99b, MP01a, MP01b, MA03b, MA03c, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, MGH09, MS06b, MV12b, MKR^{+}04, Mit05, Mit06, MS08a, MMK^{+}05, MS00b, MSG06, MMT08, MT99b, MKB^{+}09, MS00c, MPV01, MF02, MTC^{+}08, MVA01, MAFSS04, MS96c, Mon01, MMM05, MTU09, Moo01, MSK^{+}12, Mor95, MR12c, MDPQA09, MDPQA12a]. Conference [MDPQA12b, MMC05, Mor06, MK96, MBF08, MHW12, MMH07, Mos97, MNS95, MSS04, MSY00, MR11, MSX10, MSS08, MAR05b, Mue06, MOTN07, MTNL12, MC98, Müh75, MPS04, MCER10, MSUB04, MSKY10, MBC^{+}97, MV06, MA07c, MYS10, MWK^{+}09, MMM11, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, MV07, Mur99, MBH^{+}11, MZ06b, Nac01, Nad99, NWL06, NNS03, NPPDG06, NNPW12, NR07, NP09, NP11a, NP12a, NZZ11, Nao04, Nao07, NNS06a, NNS06b, NR99, NO11, ND09, NJ10, NP12b, NDF94, NR98, NHJ04a, NHJ04b, NHJ04c, NT06, NKPH08, NSM07, NKLC06, Ngo09, NKN^{+}05, Ngu06, NBGA08, NKC09, NLŚ10a, NLŚ10b, NP10, NKJ11a, NKJ11b, NHJ12a, NHJ12b, NS87, NL09, NCBT04, NT98, Nie99, NS00, NE02, NZ03, NKM^{+}09, NdG04, NL99]. Conference [NWHR06, NV01a, NRWI09, NV01b, Nie03, Nie05, NRR12, NBK04, NW10, NS98, NCJK08, NQL06, Nip98, NK99, NPS06, NP11b, NLS08, Niv98, NT03, Nob12, NABK12, NBC^{+}06, Nor04, NR08, NS90, NG10, NQR09, NBM09, Nyb98, NSS03, O’B12, OSR^{+}02, OVG^{+}08, ONC^{+}11, ODP^{+}07, Ode04, OT11, Ohl93, OiDP98, OKHH^{+}08, OKS02, OS07, OYM07, Oka00, Oka04, OW07, OA07, OP97, Oli07, Olt95, ORB^{+}09, OS03a, Ong05, Ong10, Ong11, Oqu07, OYA04, OT07, OWPY03, OALZ^{+}08, OD12a, OKBMB10, ODV08, OF15, OV09, OSRS07, OZ09, OZ11, PR01, PR08, Pag98, PRV11, PHR09, Pai09, PPP07, PM04a, PM04b, PM04c, PM04d, PS02a, PKM04, PHJ03b, PHJ03a, PGM98, Pal00a, Pal03, PVW10, PDH99, PDH01, PDHS03, PAD^{+}08a, PCH^{+}06, PAD^{+}06, PZE^{+}08, PCN10a, PCN10b, PCN10c, PCN12a, PCN12b]. Conference [PCN12c, PCK^{+}12, PKM^{+}10, PDSD10, PSDS11a, PSDS11b, PDSN12, PP12a, PH04, Pap95, PH11a, PA09, PBK^{+}10, PSST07, PV00, PCA^{+}09, PZYS12, Par09, PSS^{+}10a, PSPH^{+}10a, PSPH^{+}10b, PdRM^{+}12, Pat03, Pat11, PFD^{+}10, PTB99, PAD08b, PDMFF12, PMT09, PMT10, PRK12, PYLW08, PMA^{+}07, PV02, PLC11, PCPS03, PF06, PF08, PF10, PFM12, PR03b, PI05, Per06, Per07a, Per07b, Per08, PS08a, Per09a, Per09b, Per10a, Per11a, Per11b, Per12a, Per12b, PT98, Per10b, PT92, PT08b, PCS^{+}07, PS04b, PT97, PW03, PVTG07, PSM10, PPS03, PRH10a, PRH10b, PVP05, PMKG07, PY05b, Pez03, PHH10, PB09, Pfe94, PS03, Pfe06, Pfe07, Pfi01, Pic01, PH12, PMWO02, PSN95, PSNS99, PGD07, Pie08, PZ09, Pie10, PBK03, PK12, PH06, PP03, PA03b, PDG92, PS11a, PRJ95]. Conference [PRG09, PRG10, PRG11, PK00, PHOK10, PVV10, PP11, PMC06, Poi06, PJ12, Pol86, Pol00, PS98b, PS09b, PLK^{+}12, Por98, PR00, PS07b, PG09, PISW05, PT91a, PLI08, Pre00, Pre02, Pre09, PLMV12, Pri06, PPM09, PCA02, PPH07, PR98b, PBW05, Pro11, PW07, PW08, PSS98, PPF^{+}00, PP04, Pup99, PS09c, QQ11, QOZ01, QGZ03, QIW07, QMW09, QSWL11, QLXL07, QLK^{+}10, QLK^{+}11, QBA01, QSBC98, QBY^{+}12, QS00, Rab10, RKS89, RWSS12, RMDS10, Raj09, RFBW12, RW11, RR08, RLL04, RRR99, RCA90, RS06, RS09a, RR12, RA12, RM02, RAW07, RD01, RCJ^{+}08, RT07, Rau11, RMFG06, Rau04b, RLKHL12, Rea01, RV10, RDPMA^{+}11a, RDPMA^{+}11b, RD12a, RK06, Rei09, RW06, RRR01, Ret97, Reu99, Reu01, RMS^{+}05, RL06, RN11, Ric02, RRM06, Ric98, RYG11, RMTW11]. Conference [RBTD04, RPBP06, Rob11, Roc07, RCLE05, RS09b, Rog11, RLT^{+}08, RBC93, RBW95, RV11, RCS10, RJS^{+}08, REPH06, RS03, RT10, Ros03, RT09, RBGG06, RS01, Rov98, RO00b, Roy05, RS09c, RJL^{+}08, RDK09, RG11, RB00, RGLA09, RBB^{+}06, Rus00, RSTZ04, RTZŻ06, RTZZ08, RST^{+}10b, RST^{+}10a, RKS^{+}12a, RKS^{+}12b, RKNV09, RSK03, RS09d, RSW12b, SS07a, SPBP08, Sad13, SP09, SNČ^{+}12a, SNS03, SNY06, SN08, SNC12b, SO07, Sah13, SKS02, SCH^{+}09a, SM11a, SS11a, Sak91, Sak01, SGPP06, SL12, SHJ09, SSO^{+}11, SYNC11, SYMA09, San92, SZR^{+}08, SPCG07, Sd98, SVS02, Sar03, Sas05, SNE08, SINK08, Sat97, ShK06, SPH07, SCKR10a, SCKR10b, SMB03, SB95, SHØ06, SSRA98, SH08, SDG^{+}07, SKMT04, SF74, SUW02, Sch11, STL11, SCM88, Sch06, Sch09b, Sch12]. Conference [Sch74, SR07, SEG09, SF11, SB07, SMH^{+}12, SDR^{+}00, Sch10b, SW03a, SJW^{+}12, Sch07, SSN08, SS09b, SS06a, Sco02, Seb03, SVS08, Sei07, SHGB05, SHW09, SP08, SZC^{+}03, SO09, SC09a, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, Sex09, SEO^{+}09, SMP07, SHX^{+}06, SW09b, ST02, SW08, Sha86, STS04, SCL^{+}06, SLL^{+}07, SYY^{+}08, SWJX12, SdDK91, SSD^{+}08, SvADS07a, SvADS07b, SvADS07c, SvADS07d, SKSA12, SILS05, Sho84, Sho05, SLC01, Shu07, Shu09, Shu11a, Shu11b, Shv06, SU97, STA^{+}11, SCR06, SAG10a, SAG10b, Sie75, Sie86, SKB10, SFN^{+}11, Sil01, SCP^{+}12, SMK01, SSAP05a, SSAP05b, Sio10, SMPU98, SEL09, SB12a, Sir12, Ska01, Sle92, SYP^{+}05, SA12, SBH98, Slo99, STDH02c, STDH02a, STDH02b, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, SCH04c, SHP^{+}05, Sma97, Sma05, Sma08, Smi12, Smo97, Smo00, SLSW12]. Conference [SKP02, SKP04, SKP06, SHKP08, SHKP10, SHKP12, SCZ10, SL99b, SP93, Son00, SLLS03, SEN07, SEN08, SQL10, SC09b, SMK02, SLKL09, SH07, SS08b, SR11, SXC05, SRY07, SB12b, SS06c, SSU12, SKCK07, SHL08, SA97, SDH03, ST12c, Ste98, SL03, SH10, SSB90, SDK^{+}08, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, STTP02, Ste99, SWB03, SS09d, Sti90, SRM^{+}11, SD10, Sti94, SWMM06, SS10, Str96, SM01, Stu02, Stu08, ŠPTM00, SDM05, SCB01, Stü09, SRL12a, SRL12b, SRL12c, Sub05, SHRM06, SLCT11, SNSR06, Sur02, SLM07, SMS12, SHYEF08, SHUH08, SC00, SBH^{+}04, Syv02, SV12, SKN05, SHŚ^{+}07, SKR^{+}10, SH11, SSZ12, TR12, Tah09, TMLL09, TOOO07, TWQ^{+}12, TDWG11, TBE08, TMG10, TMdB11, TMS12]. Conference [TSG00, TLF01, TST10a, TST10b, TSCW11a, TSCW11b, TCHB12b, TCHB12a, TSJ12a, TSJ12b, TWYhL01, TLZ^{+}08, TKCW11a, TKCW11b, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, TG04, TC99, TN03, TT10, TY11, TdLdOQ08, TTM10, TDS94, TLC00, TS10, Tha96, TBC07, TKCH09, TKSN11, TCYZ12, TPCB12, Tho99, Tho06, TG12, TBH03, Tis02, TN09, Tiu00, TT06a, TQYX11, TLJ06, TP94, TBF^{+}09, TO08, Tom91, Tom92, TNP^{+}07, TN04, TNM05, TNVDF06, TNY07, TN08, TNI09, TND10, TNYL11, TCJ10, TTF07, TG10, TL02, Tra04a, Tra04b, TGAH11, TKHS05, Tre09, Tri98, TMN09, TGM97, TCLH10, TOL08, TdRM^{+}09, Tsu08, TSKGB04, THT03, UPB07, UWB^{+}06, UMN^{+}08, UBW06, Uus06, UDF06, Vad07, VKV00, Val94, VGP08, VTK^{+}12, VJ11, VY93, VPBŠ03, Van02, VW07, VGM^{+}09, VM99, VM01, VW05, VV03]. Conference [Vau06, Vau08, VFC08, VRHJ09a, VRHJ09b, VA95, VG12, Ver08, Ver00a, Ver00b, VU09, Vet11, VMBMN04, VZM^{+}10, VKMT11, VDÉ^{+}12b, VDÉ^{+}12a, VW08, VV07, VD12, Vit08, Vit10, VSH11, VGB11, VS02, Vog01, Vog88, Vol93, Vor92b, Vor02, Vor08, VLY09, VWY01, Wad87, WAS06, WT96, WRP07, Wag08, WSS04, Wal99a, Wal01, Wal04b, WYL01, Wan01, Wan03, Wan04, WPV04, WJ05, Wan05a, WJS^{+}06, WPSY06, WLC^{+}06, WmCL07, WPR07, WPR08, WLGB^{+}08, WFZL09, WGMP09, WDGL10a, WDGL10b, WGLL10, WR11, WZD11, WLO^{+}11, WS12, WZJX12, WB00, WZ03, WSTI07, WSTI08, WY99, WW03, Wat00, WY04, WK02, WR07, Wel01, WC98, WW02, WLP^{+}09, WMS04, WL04, Wha03, Whi00, WCK11, WTP^{+}06, Wie99, WP10, WBW94, Wil01a, Wil99, WM04, WM03, Wim03, Wim04, WSGA06, WSG07, WSF08, WSJT09, WCJS10]. Conference [WFB12, WL03, Win11, Win09, WDKK07, WZA10, WN96, WZ08, WPD04, WPZA11, WZ11, Wol95, WKS^{+}10, Won01, WK06, WMB10b, WMB10a, WANL08, Wri03, WKK98, WCGB05, WYX08, WLRL08, WD10a, WD10b, WKD08, WDPW02, WDPW04, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, XCd03, XCHZ11b, XCHZ11a, XSA^{+}12a, XSA^{+}12b, X^{+}12a, X^{+}12b, XYM^{+}07, XKGS12, XXX^{+}09, XBS^{+}10, XCR11, XHXL08, XLHX08, XL11, XYZU11, XBM12, YBM87, YKKZ07a, YKKZ07b, YNB08, Yak00, Yak02, Yak04, YN06, Yam08, YGGJ04, YEY04, YJMU06, YW06, YDW08, YHR^{+}09a, YHR^{+}09b, YYI^{+}09, YMHH10, YGE12, YKF97, YLW^{+}07, YSP^{+}10, YRWS11, YYS^{+}12, YT99, Yas00, YI12, YLWY06, YX04, YAF^{+}02, YTC^{+}07, YWRS11, YCB12, YM01, Yor08, YMY^{+}10, YP01, YLLZ04, YKL06, YGL^{+}10, YLC^{+}10, YKU11a, YKU11b, YYH^{+}12, YZCW07, Yun02, YDKM06, YLL09, ZBW94, ZYRP10a, ZYRP10b, ZRP^{+}09, ZRP^{+}11a, ZRP^{+}11b, Zam08]. Conference [Zan00, ZEL04, ZTTY12, ZBRL12, ŽL08, ZPT^{+}06, ZiTM10a, ZiTM10b, ZiTM10c, ZJ04a, ZZ04a, ZGY04, ZJ04b, ZJ05, ZTY^{+}05, ZS07, ZOCJ07, Zha07, ZYBX08, ZPTI09, ZZP^{+}10, ZCDT10, ZS10, ZO10, ZHLW12, Zhe02, ZZ99, Zho01, ZLLC09, ZCGH11, ZYH03, ZNCF03, ZZO03, ZSP^{+}04, ZYB06, ZLS^{+}06, ZLY07b, ZW09, ZYDL09, ZY10, ZZLL10, ZYRH06, ZF05, ZY01, ZBH12, ZKZ00, Zim92, ZVU99, Zit01, Zuc03, Zui99, ŻR99, dA09, dBBC02, dCVK07, dFEN04, dGEK11, dKIM^{+}07, dLZ12, dMLG03, dMS09, dRF01, dRS01, dRWK^{+}10, dSADM07a, dSADM07b, dSDL04, de 98, dS02, dS11a, dCHM12, dPMA98, dlBP08, tTVH^{+}12, ter97, vBC08, vB97, vDKS11, vEGW09, vEGSW11, vHV08, vHC08, vHH09, vH83, vF91, vLWH^{+}00, vLIvdH^{+}07, vLMP^{+}10, vO04, vSN01, vW97, vSH06, vdBHOS09, vdBGG10, vdBV01, vdHI99, vdHHsHD06]. Conference [vdHCD07, vdHXMW08, vdHS10, vdHIP11, vdHP12, vdAB03, vdAHW03, vdMvdT08, vdV84, vGH96, ABK05, AN95, ANS^{+}09a, ANS^{+}09b, ADB^{+}05, AH96, ADW05, ADL05, Ang97, AM04, ACL^{+}04, Aya95, AB97, BV05a, BPSP06, BM05b, BUG93, BMH05, BG91b, Bec05, BIRT05, BCT06, BMP90, Bha95, BMW93, BN91b, BBP93, Bla05, BTPPS05, BGU94, BGU95, BC05, BMM07, Bod96, BBD97, BCP94, BP05, BP04b, Boy95, BG05, Bra75, BCKT05, BL06c, BW05b, Bro92, Bro94, BRR97, BvADS08c, BC04b, BR96b, BT06, yCW96, yCCZ07, Cal86, CCP96, CDP^{+}05, CE95, CFB^{+}05, Cea76a, Cea76b, CTT93, CV96, Chi79, CD05, CAZ05, CM05a, CN05, CLT05, CLP05, CFP94, CP05, Cou93a, CK82, CM05b, DKP05, DWSK05, Dar97, DG96, DV05, DN04, DR05, DE05, DIMV11, Del97a, Del97b, DL04a]. conference [DJ90, Des86, Des94, DWML05, Deu81, DEM96, DPA05, DF96, Dil94, Din04, DRS05, DOSY05, DLW06, DL95b, DL81, DG05b, EFM04, Ehr04, EL04, Esi93, ER05, EKL^{+}05, FLS06, FSdlPT11b, FHRR05, For88, Fra05b, FP05, FK95a, FKKB^{+}05, FG04b, FTF88, FTF89, GO94, GG05a, GP05, GHM05, GG05b, Gan96, GL97, Gel05, GWS05, GM04a, Gib89, GHSW04, GL05, God05b, GPPW08, GT93, Gru97, Gut94, HZ05, HOQ97, HR96, HHM97, H^{+}05, HK94, HL04, HSSY05, HIS05, HS97, HML97, Hsi95, HM04, Hug91, HM05b, JPBT05, JP05a, JKB04, Jen92, JE06, JL97, JRJ05, Joh97, JP94, JS84, Jou94, JK05a, JK05b, Kal97, KC05, Kar94, KL05, KTC^{+}05, KD05, KHJ05d, Kil05, Kim05a, KSW05, KW90, KL92, KKKN05, Kit88, Kob96, Koh06a, KK95, KA05]. conference [LLC^{+}05, LA94, LWD05, LVW05, LCHS96, LSS^{+}04, LDTU04, LMV95, LMW06, LS89, LR94, LSL04, LB78, LQO04, LD05b, LD05a, LFL05, LG05, Luo05, Mad89, Mah85, Mai90, Mal95, Mal97b, Mal05, MNT95, MR97a, MBS05a, MJR04, MMT97, MMP05, MSW05, MXZ05, Mel86, Mic90, MR97b, Mol95, MR95, Mor89a, MSH05, Nak92, Nie93, Nor86, Nor87, NK88, Nor90, eOP05, PBB05, PDML11, PGA^{+}05, PC05, PF05, PY05a, Per05, PS96b, QMLW05, RG05, RS97a, RMS04, RCT05, Rei95, Reu97, RT95, RAM05, RV05, RSvdP05, Sam76, SP90, Sel86, SLB05, Shy92, Shy93, SJT05, SBG04, SC96, SZ05, SW75, STLK05, SvASD05a, SvASD05b, SvASD05c, SV05, SA04, SC05, TPB79, TLC06, TTP05, Thi94, Thi95, TT05, Urz05, Van05, Van89, VPM97, VH04, Vit95, VBCBS05, Vor92a]. conference [Vor93, VP04, Wad86, Wal04a, WCO05a, WCO05b, WCO05c, Wan05b, W^{+}12, Wei79, WTGA05, WGD05, YZZ^{+}05, YAL^{+}05, YRMD05, YBL^{+}04, ZJ06, ZOM05, ZLDB05, vB05, vdABCC05, von96]. Conferences [AL11a, ABD^{+}09, AI06, Ama08a, BH06b, Bar11, BPR05, Bir12, Bod05, BKL^{+}12c, BBK^{+}11, Cas09a, Cer05, CW09, Cle99, De 07b, Deg03, DG12, Dro08, Due04, DBB^{+}08c, DL07, EM10, Est98, FI08, Fin99, FK12a, GH03, GO11, Gor03, Gor10, GS00b, GH07b, GHP^{+}08, Gup10, HZ05, Han98, Hed03, Hen08, HP06, Hof11, Hor02, Hus01, Jäh99, JP04a, KS02c, hKA10, Kno11, Kos98, KP09, KO07b, Le 02, LFP^{+}09, Mai00, MY01, MZ98, MTP^{+}02, MTS^{+}03, MT05a, MT06a, MT06b, MT07b, MT07c, MT08a, MT08b, MDH09a, MDH09b, MDH10a, MDH10b, MDH^{+}11b, MDH^{+}11c, MZ06b, NE02, Niv98, O’B12, Ong10, PS07a, Pez03, RR08, RT10, Sag05, SHN07, San01, Sas05, Sch04b, Sei07, Sei12, Ses06, Ste98, Swi99, Tho99, Wal04b, Wat00, WK02]. Conferences [WMS04, Wil01a, dA09, dGN12, dLZ12, dMS09, LY04]. Confidence [Mei08]. configuration [Con97, Est95, Est99, Mag98, Som96, WvdH03]. confrence [HM04]. Congress [AMGJ12, AGS91, BM05a, BP07, Bau99, BdBK03, BvDMST11, BC10b, CT03, CD09, Esp01, FvdHJT10, HSP^{+}12, IT06, KGV^{+}09, LM00, Len97, LABM^{+}12, MTOK06, MDY07, MCA^{+}07, PPCS12, RMK07, RSK08, SK11, SRS03, SL05, Tor93, WWD99a, WWD99b, ZYW08, STC04]. Conjunction [CPR^{+}10, CKP^{+}12, DFGN12, hKAS^{+}11, LMCH12, LSIT11, MDY^{+}10, MDHB11, RKM^{+}12, RKS^{+}12c, SWSY11, WYSS10, YLS^{+}12, YC11, YSL12, YHV12, Rol98, Rol99]. CONLAN [PBB^{+}83]. Connected [BKM06, HL05]. Connecticut [Ban93]. connection [Bau98]. Connectionism [MP01b]. Connectionist [MP01a, Van94, RS96]. connotational [Roy87]. CONPAR [BCRT92, BV94, Bur90, Han81, HHJ^{+}86]. Consequences [GGHS07]. Considering [KI01]. Consolidated [TDBP01]. Consortium [GKMN10, MTH06a, MTH06b, MTH07a, MTH07b]. Constants [MO06]. constrained [FP90, PR87, JFM96]. Constraint [AB11, AFR^{+}04, ABFR07, BS06a, Ben06, Bes07, Bis04, BJN03, BRM06, BJ01c, CW02, Coh10, Dec00, Esc10, FRS08, FPFR05, Gae96, Gen09, HCF06, Kuc11, LO11, Lee11, LMT10, MP98, Mar11c, Nar01a, O’S03b, OFR09, PT08b, RR04a, Ros03, Sch02, Stu08, Van02, VW07, Wal01, vHH09, BM05b, Bor94, CKPB07, DHM00, Fre96, GH92, Jaf99, JNS05b, Mil12, MR95, Smo97, Wal04a, vB05, BRM06, CSV05, Gae96, KR04b, KW96, Mey95, Nar01a, Pod95, SF08a]. Constraint-Based [Nar01a, GH92, BRM06, Nar01a]. constraint-related [CKPB07]. Constraints [AFR^{+}04, ABFR07, CKV08, FRS08, FPFR05, Gae96, HCF06, LO11, OFR09, Apt00, O’S03b, CMT01, Jou94]. Construction [AL11a, AG09, APM08, Bod05, BM02a, BDK10b, Due04, EM10, FK12a, GN12c, GH07b, Gup10, Hen08, HP06, JP04a, Kno11, KP09, Koz04, KO07b, MZ06b, O’B12, RR08, TO08, Uus06, dMS09, BO00a, BCG02, BE74, BE76, BB79, BMW93, Bri95, Bri97, BJ95, Cle99, Fri94, GH03, GS00b, Gyi96, HZ05, Hed03, Hor02, Jäh99, Jeu98, KP92, KS02c, Kos98, MS96a, MY01, Mol95, Ste12, Ste98, Van89, Vog92, Wat00, Wil01a, BBH^{+}05, BGH06a]. constructions [GB77]. Constructive [SH12b]. Constructivity [MO92b]. Contemporary [SSW92]. Content [BHMS07, CMSM^{+}10, CL04, CvSW04, DJW04, GMS03, GKS12, GJTS06, MGSM12, SLZH07, AGL06, NK99, CL04]. Content-Based [CMSM^{+}10, MGSM12]. Contents [YAKMJ11]. CONTESSA [KW96]. Contest [ÜÇA10]. Contests [ÜÇA10]. Context [Akm01, BMPM09, BCRB^{+}11, BSB^{+}99, CQSS09, CR05b, DJNR06, DKN^{+}11, DKLT05, EEM^{+}11, FFL^{+}12, HLMW06, HKS06, HSS07, KRRBV07, KFLS07, LNSS11, LKK10, MMBND07, RLT^{+}08, RBSL06, RFBD09, Ste11d, Bla03a, SLP05, Akm01, Bla03a, Nij80]. Context-Aware [EEM^{+}11]. Context-Awareness [HKS06, HSS07, SLP05]. Context-free [Nij80]. CONTEXT’99 [BSB^{+}99]. Contexts [DS08]. contour [FMdGC99]. Contraint [ABFR07, BJP12]. Contributing [KM11]. Contributions [PT91a, GT87, Cav85, GOW03, GT90, KK87, WWC02, YBM87]. Control [AK10, AC09, AP04a, Bee02, BBB07b, CT07, DF02, EM08, HT06a, MT09c, NQR09, Pas02, Rei03, RHD08, SHS10, Bau00, BHGP02, BV92, Cas81, CW94b, DSV01, DER94, HS98a, LK00, MP03a, Man75, MT05b, MSS05b, Nai86, PS97b, Per05, Sch91, TG02, VvS99, ABL96, AHS96]. Controlled [Fuc10, Fuc10, KF12, RF12]. Converged [CMS11, MDY07, PAD08b]. Convergence [DS08, KT06, OSC^{+}08, SKCK07, Del00, Kim05a, LHŚ11, LHKŚ12]. Convergent [Kou11]. Conversation [Hug03]. COOP [AC09, CT07, NQR09]. Cooperation [DL81, KK11c, LB78, Sam76, Had96, Lub95, PS97b]. Cooperative [HTW02, JPXS04a, JPXS04b, LSDN04a, LSDN04b, SCL^{+}06, SLL^{+}07, SYY^{+}08, ZF05, Bui87, HRR93, MBC^{+}97, SLB05, SLF91, Bat01, BB01b, DMV01, ES00b, KK97, KW98, KSW99, KK00, KZ01, KOS02, KOOL03, KOK04, KRP06, KHPS07, KPP08, Lib07, Luo04, Luo05, Luo06, Luo07, Luo08, Luo09, Luo10, Luo11, Luo12, SKB98, Str99]. CoopIS [MT05a, MT06a, MT06b, MT07b, MT07c, MT08a, MT08b, MDH09a, MDH09b, MDH10a, MDH10b, MDH^{+}11b, MDH^{+}11c, MTS^{+}03, MT04d, Bat01, ES00b, MTP^{+}02, MT04c]. Coordinating [von92]. COORDINATION [CA10, DR11, FV09, JP05a, LZ08, MV07, Sir12, CNY95, CW06, Sch01b, Pad99, AT02c, BPD06, CH96, CW99, CW06, CA10, CN98, CvRVSN12, DR11, DFM04, DFPV11, FV09, GL97, HMBD09, JP05a, LZ08, MV07, NVSB^{+}07, PAV^{+}10, PR00, SPON08, Sir12]. Copenhagen [ADW05, AMCL02, Arn90, BG10a, BC95, BJ10, BG12a, Bjo80, BL02b, CF09a, EF02, EL02a, FS09a, GJ86, HMT99, HS02, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, Hut96, Jon90, KMYJ08, LNS06a, LNS06b, LPK05, MPB05, MFG12, NSFB10, PHOK10, PBW05, RR01, Stu02, Tis02, TBJ02, TT05, Tou94, Vor02, WTGA05, YKMJ09, YAKMJ10, dGN12]. Coping [Ste07d]. CORBA [SLPM96]. Cordoba [GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, CFVCF^{+}12]. Core [CZG^{+}08, Sto04]. CoreGRID [LMSS07, GVT^{+}11]. Corfu [ABK^{+}09, CMP07, DFM11, Nie99, PDK^{+}10, PCG^{+}10, Rei88, Spi95, TZS00]. Cork [BR04, BS06a, BBK^{+}11, O’S03b, RS09b]. Coroutines [Mar80]. Corpora [KMPH09]. Correct [LdRV94, BJ95, BP05, CE95, ELLP12, GT03, LK99, MM01, MP93, SO99]. Correcting [BS01b, BL07b, BAH09, FILP06, BC88, Cal86, CMM93, CGM95, Fos99, FHP03, Gur04, HP89, MMR91, Mor89a, MM97, Pol86, Sak91]. correction [Pet80]. Correctness [DHS10, Zwi89, dBdRR90]. Corrigendum [TDB^{+}06, TDBP01]. cortical [ÉEMS04]. Coruña [DMV12]. Corvallis [BRST08]. COSADE [SH12b]. Cosenza [Fle02]. COSIT [CM05a, EGMW11, HCDL09, WDKK07, FM99, KWT03, Mon01]. COSIT’93 [FC93]. COSIT’95 [FK95a]. COSIT’97 [HF97]. COSS [RBsC^{+}06]. COST [EFZKM07, EBAH08, EV09, EHMM09, ECV^{+}10, EVV^{+}11, EEM^{+}11, FOGE^{+}09, HS12a, SSKR02, TSE^{+}09, VEG^{+}11, VDD^{+}11, dSOSR06, MWD06, CRS00, DD97, Hut94, Hut96, KS03b, RMV96, Smi01, Smi03, SCKR03, VDPD96, dS02, dSOSR04]. COST264 [CH01, RF99a]. Costa [dVGR04]. COTS [DG02, EW03, FP05, KP04]. COTS-based [EW03, FP05, DG02, KP04]. Cottbus [Tha96]. County [DF06, DS06]. Course [Bra80, Enc80, PR98a, And95, BRR87a, BRR87b, FHM82, KBD87, LMŠT01, MSH85, Nos88, PGV93, TV87, ASP85, BE74, Bau75, BE76, BSG84, BJ86b, CJ95, GSB78, LSDP81, LPSD83]. Courseware [Ull08]. Coventry [BND01, GM89, JLY03, SCL^{+}06]. covers [Nij80]. Covilhã [BCD05]. Cozumel [Tsu08]. CP [Ben06, Bes07, Coh10, Gen09, Lee11, Mil12, Stu08, Wal04a, vB05, Dec00, Fre96, Gae96, MR95, Ros03, Smo97, Van02, Wal01]. CP98 [MP98]. CP’99 [Jaf99]. CPAIOR [AB11, BM05b, BS06a, BJP12, LMT10, PT08b, VW07, vHH09]. CPM [AP10, AH97a, ACP05, FL08, GM11a, KS12a, KU09, LV06, MZ07a, AL01, Apo93, AT02b, BYCC03, CG94, FC98, GU95, GS00a, HM96, PC99, SMD04]. CPP [HM12b, JS11b]. CQRE [Bau99]. Cracow [BDW97, DKN02, Fra05b]. Crakow [PS96a]. Crans [BCB97]. Crans-Montana [BCB97]. Cray [RR89]. Creating [DKW07, SKS^{+}09, TP97]. Creation [XCR11, dMLG03]. Creativity [TKSN11]. Crema [BPT06]. CREST [Enc80]. Crete [AB12, APS06, AGS^{+}11b, AGS^{+}11a, AZ11, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, BBP^{+}12a, BBP^{+}12b, BCP^{+}04, BDFS04, CCPP12, CCK95, CMV96, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, DKW07, GPE05, GS06, Kut12a, Kut12b, LW95, LM06b, LMT^{+}04, LDC^{+}09, MSS12, MS10b, MS10c, MS12b, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, NT06, NS98, NT03, OSvL01, PR11, SMBQ07, SCP^{+}12, Thi10, TP09]. Creusot [Col02]. Cristian [DKN12]. Criteria [Nes95, Bui87]. Criterion [EFG^{+}09, ODP^{+}07, TDWG11, CAZ05, Fon03, Zit01]. Critical [ACJ09, BRS^{+}07, CF09b, Gie10, KR10b, LM08a, SS11b, SP12, Koy92, RBS93, Lop06, LH08, LSW12, RB10, SG09, XW11]. CRITIS [Lop06, LH08, RB10, SG09, XW11]. CRIWG [BAdVR08, CBF09, DZGS06, FLS05, HOC07, HHJZ12, KHL10, VGB11, FD03, HP02a, dVGR04]. CRN [CGMP11]. Croatia [BHM^{+}04, DM06, JKN^{+}12, LGTB03, LGBT03, LHJ08a, LHJ08b, LHJ08c]. Croisic [DMR01]. Cross [AFP^{+}10, CCM^{+}04, FGK^{+}11, Pet01, PBGK03, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, QBY^{+}12, TQYX11, vdBK05, PBGK02, PGBK04, EV09, Pet01]. Cross-cultural [vdBK05]. Cross-Domain [QBY^{+}12]. Cross-Fertilization [CCM^{+}04]. Cross-Language [AFP^{+}10, FGK^{+}11, Pet01, PBGK03, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, PBGK02, PGBK04, Pet01]. Cross-Modal [EV09]. Crosscultural [Ish10]. crosscutting [YM01]. Crossroads [AM03]. Cruces [PN93]. Cruz [Ell95, NNPW12]. CRW [DTF98]. CRYPTO [BC85, Bra90, Dwo06, Fei91, Fra04, Hal09, Men07, Rab10, Rog11, SNC12b, Sho05, Wag08, Wil86, Bel00, Bon03, Bri93, Cop95, Des94, Gol90, Kal97, Kil01, Kob96, Kra98, MZ98, MV91, Odl87, Pom88, Sti94, Wie99, Yun02]. Cryptographer [Abe06, Dun12, Fis09, Joy03, Kia11, Mal08, Nac01, Oka04, Pie10, Poi06, Pre02, Men05]. Cryptographic [BCI85, Bih03, CC04, CP88, Cra05, Dam91, De 95, Fum97, Gil10, GQ95, Gün88a, Hel94, Jou09, Knu02, Mau96, Nao07, OF15, Pat11, Pfi01, PJ12, Pre00, QV90, Rue93, SZ93, Sma08, Vau06, Nyb98, Pic86, Ste99, CG09, GM06b, JQ04, KKP02, KP99, KP01b, KNP01, MS10a, OR08, PV07, PT11, PS12c, RS05, WKP03]. Cryptography [APFV09, AKS09, BDZ04, BSZ12, BGKY08, Bet83, BY07, BGS11, BDC12, BKNS06, BD08, Can08, CFGN11, CM09a, CSN14, Cra08, Cra12, DDS12, Dan12, Dav91, DY06a, Des02, DR06, DD07, DG09b, FBM12, GP08a, GD10, Gol11b, HR06, HH04, Hir98, IKY05, Ish11, JRSN09, JYZ04, JT09, JMO10, Jue04, KY07a, KM08b, Ker12, LT11b, MZ04, Mic10, MV12a, Nao04, NP10, OW07, ODV08, Par09, PR98a, PT06a, Rei09, Sad13, Sen10, SW09b, Sio10, SCD^{+}10, Syv02, TOOO07, Tsu08, Vad07, VY01, VD12, Yan11a, Ytr06, YDKM06, ZYH03, ZYB06, ZY10, vDKS11, AMW07, BFS92, Bla03b, Fra01, Fra99, Gab13, Gre94, HA00, Hir97, IZ98, IZ99, IZ00, JY14, Kil05, Kim01a, Lin14, NP02, NH03, PY05a, PGV93, Sah13, ST01, TM99, Vau05, Wri03, Bet83]. Cryptography [Boy95, Che11, Dar97, DG96, Gal07, Hon01, Nac12, Par09, Pat03, Sil01, Sma05, Wal99a]. Cryptology [AB10, Abe06, Abe10, BYLJ11, BL06a, Bel00, BL10a, BC11a, BC85, Bon03, Bri93, CC04, CLL^{+}09, CGL^{+}11, CRD08, Dun12, Dwo06, FLY06, Fis09, Fra04, GN12a, GG10, Hal09, HN12, JM03, Kia11, Kim12, Knu02, Kur07, LC06, LYL11, Lai03, LC09, LH10, LW11, LL04, LYL06, LSW08, MMV06, Mal08, Mat09, MZ98, MV12b, NR07, Nao07, Neu04, Ngu06, NP11b, OiDP98, Oka00, Oka04, OF15, Pat11, PYLW08, Pie08, Pie10, Poi06, Pre09, Rab10, RD01, RL06, RN11, Rog11, RO00b, RS09c, SNC12b, Sma08, SRY07, Sti94, Vau06, Vau08, Wag08, WS12, WK06, YLL09, Zhe02, AL12, BCI85, Bih03, Boy01, Bra88, Bra90, CV04b, CP88, CW94b, Cop95, Cra05, Dam91, Dam99a, Dav91, DV05, De 95, Des94, Fei91, Fum97, Gil10, Gol90, GQ95, Gün88a, Hel94]. cryptology [IRM93, Jou09, Joy03, Kal97, Kil01, KM96, Kim02, Kob96, Kra98, LOX99, LL03, Lee04, Mau96, MV91, MS02b, Men05, Men07, Nac01, Nyb98, Odl87, PC05, Pfi01, Pic86, PSN95, PJ12, Pom88, Pre00, Pre02, QV90, Roy05, Rue93, SP90, SZ93, Sho05, Son00, Ste99, Wie99, Wil86, Won01, Yun02, BLO^{+}07, DWML05, FHW08, GMO09, HWG10, LTW11, PMC06]. Cryptosystems [IZ00, Kim01a, NP02]. CS [LBP00]. CSCLP [AFR^{+}04, ABFR07, FRS08, FPFR05, HCF06, LO11, OFR09]. CSCW [Lub95]. CSCWD [SLB05, SCL^{+}06, SLL^{+}07, SYY^{+}08]. CSE [Ehr85a]. CSEE [DH94, Ibr95]. CSFR [MSZ90]. CSL [DV10, DH07b, Ési06, GK09, KM08a, Ong05, BM03a, BBR88, BBR89, BBR90, Bor91, Bor92, Bor93, BGM94, Bra02, Bun96, CS00b, FR99a, Fri01, GGS99, MT04a, NT98, PT95]. CSL’96 [vB97]. CSLP [CSV05]. CSP [AJS05]. CSR [AM10, DVV07, FMRW09, HRSS08, HKLP12, KV11, GHH06b]. CSS [XLKZ12]. CT [Abe06, BGPR10, Dun12, Fis09, JW05, Kia11, Mal08, Men05, Pie10, Poi06, BND01, Joy03, Nac01, Oka04, Pre02]. CT-RSA [Abe06, Dun12, Fis09, Kia11, Mal08, Men05, Pie10, Poi06, Joy03, Nac01, Oka04, Pre02]. CTCS [MR97b, PRJ95]. CTIC [FFL^{+}12]. CTRS [DL95a, KO91]. CTRS-94 [DL95a]. Cuba [NV01b, RSK08, SRS03, SL05]. Cube [PPPM10]. cubic [MP96]. Cuernavaca [Whi00]. Cultural [CFVCF^{+}12, FdCLN^{+}11, IFL^{+}12, SYNC11, XCR11, YGE12, vdBK05]. Culture [Ayk07b, Ish10]. Cumberland [Die03]. Cup [AK99, BLNG10, BCT02c, BJNT06, IMWZ09, KLR03, Kit98, LSST07, NRSSV05, PBBY04, RMSS12, dSCP11, SBK01, VPK00, VROD08]. Curitabia [MB97]. Curitiba [B^{+}12a, BK96, LLO08]. Current [APSZ02, BHH^{+}04a, BSW95b, BEJvR11, BFG^{+}12, ČGH^{+}11, CGBZ08, CCNY10, CKV08, CKLM09, DO89b, DRSS01, GKB^{+}08, Gol11a, GTW02, GHH^{+}06a, GP02, HJW00, HSTU99, JKB96, MG07b, NKM^{+}09, PR01, PTB99, PJ97, Rov98, SF08a, SKR^{+}10, TSKGB04, VPBŠ03, WTP^{+}06, YYS^{+}12, vLIvdH^{+}07, vLMP^{+}10, AYB95, AY96, ABGP05, Che99b, DC01, PS98b, TDS94, TLC00, VBCBS05, ZY01, BCC07, CUdlC04, DF10, GHI^{+}06, HK12, Kum90, LMT^{+}04, MOBS06, MMG10, ŠPTM00, Wet92, dBRdR86]. Curtea [PPJRN^{+}10, Pau03]. Curtea-de-Arges [Pau03]. Curves [Ahn04, DFL^{+}10, BCC^{+}12, MP96]. Customer [KMSS02]. CV2 [CJE82]. CVAMIA [BS06b, SKK04]. CVBIA [ZLJ05]. CVM [HM12c]. CVRMed [Aya95, TGM97]. CVRMed-MRCAS [TGM97]. Cyber [FST01]. Cyber-societies [FST01]. Cybernetics [YLWY06]. cyberspace [KK00, XLKZ12]. Cyberworlds [GTSS11]. Cycle [Mah04, SK97, SZ00]. cycles [MvOvRdV05]. Cyprus [ABHT08, APPE09a, APPE09b, Apt00, AL11b, ADP11, BJ07, BZ08a, Cod01, Dik04, DKB03, EM10, Eug09, Gor10, Gup10, HK12, IR07, IFGH10, IFL^{+}12, KMMN11, MEK03, MS06a, MP09, MT05a, MTH05, MBF08, MV07, NS98, Ölv10, Ong10, PJC^{+}12, Pel07, Pet02, PR00, RT10, SEO^{+}09, VJ11, VAR12, Vit08, Wal01, ZBRL12]. Czech [BM05b, BSW95b, BTQ03, BBK12, BKL^{+}12a, BKL^{+}12b, BFG^{+}12, BTPP^{+}12, BJKK02, BJ12b, BR07c, CUDL02, CCMR10, CKMS05, DV10, De 07a, EV09, FKK04, FKM09, FPCK11, FN12b, Gie10, GSS09, GLM93, GV95, GGS99, GP02, GRP10, GDGB12, HM11a, HKR07, HS94, HKP10, HS95c, HJW00, HMP99, HK10b, HŽ07, HSH08, HdL12, HKM^{+}11, JKB96, JJP^{+}02, KMW03, KS01, KPP08, KK11d, KBČ^{+}12, Kra99, KLFK10, KNK08a, KNK08b, Kuc̆02, KK07, KP10c, LD97, LMŠT01, Mac06, MP03a, MJ04, MRŠ03, MMP03, MMV04, MMOS99, Mat01, MM03, MMP05, MM07, MM09b, Nes99, NKM^{+}09, OPT01, PM04a, PM04b, PM04c, PM04d, PTRH05, PP12a, PTB99, PJ97, PBRB08, QBY^{+}12, RRBE09, SLH04, SPK01, SKP00, SKP02, SKP04, SKP06, SHKP08, SHKP10, SHKP12, SKK04, SS06c, Ste99, ŠPTM00, Tra04a]. Czech [VPBŠ03, WH95, WTP^{+}06, ŽL08, ZBG98, ŻR99, vWN99, vLIvdH^{+}07, vLMP^{+}10, vW97]. Czechoslovakia [AM91c, Bec79, CK84, CJK88, DK90, GC81, GRW86, HK92, KJ91, MST92, Rov90]. Czestochowa [WDPW04].

D

[MSS02, Pal09, Pol01, BH77, KvG98, KK09a].

D.C [OR08]. DAARC

[Bra07, DBM09, HDBM11]. Daegu [LKK^{+}07, NKJ11a, NKJ11b, ZO10].

Daegue [KR10a]. Daejeon [FKLY08, LHŚ11, LHKŚ12]. DAGM [DNS09, FMNS06, GRK^{+}10, HSJ07, KK09a, KSH05, MF11c, PPBL12, Rig08, vG02, MK03a, RF01, RBGS04]. Dagstuhl [AFKL04, AKR03, BdBC01, BRS^{+}07, BHGP02, Ber03, BMS95, BM02a, CBN99, CRYS09, CMOZM11, CKLM09, DGT96, DFP^{+}12, Die02, DFN97, DdCF98, FS05, FS99a, Fre98, FW92, FLMS05, FLMK06, FLT07, FKLT08, GKS^{+}07, GKT09, GKL^{+}11, GT05, GSS00, Hag02, HHLR08, IPSK04, Jan92, JLST06, JLM00, KSHS01, KN91, KEM07, KHG01, KS05b, LBCO04, LS05, MBS05b, RSS06, RBS93, RHD08, RME04, SST00, SKP07, SE94]. Dahl [OKL04]. DAI [ZL96]. DAIS [EM06, EK10, FR11, GH12b, IR07, KA05, MT08c, SO09, SDH03]. d’Ajol [CCR86]. Dakar [CV10, NP11b]. Dal [BHK01, GW03, VvS99, vEGSW11]. Dalian [ABKT11, HW08a, JDAW12, LWF04, LJJG07, YWG04a, YWG04b]. Dallas [BMSW12, DG06b, FD11, Jay04, LHDD08, Mau89, SV01]. DALT [BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, LOTY05, OSV11, SSVW12, LOST04]. DAMA [Ari02]. DAMAS [TGH06]. Dame [AKL^{+}99]. DaMEN [XYZU11]. Dana [Gol11b]. d’Andratx [PF06, PF08, PF10, PFM12]. Daniele [ACG^{+}07]. Darmstadt [ABH^{+}10, BBC04, Car03, CS01a, Enc80, FR09, FBM12, GM00, GSH^{+}04, GMI06, GMUW12, GRK^{+}10, HG96, KS09b, KGS08, PBGK02, SDG^{+}07, SH12b, Sen10, SW97, SKB98, WPZA11]. Darwin [KKS11a, KKS11b]. DAS’98 [wLN99]. d’Ascq [GLM^{+}08]. DASFAA [CLLD09, HKP08, KILW10a, KILW10b, KKMN07, LTW06, gLPZ^{+}12a, gLPZ^{+}12b, ŠPTM00, XYZU11, YMY^{+}10, YKU11a, YKU11b, YYH^{+}12, ZOM05, ZYDL09, LLWL04]. Dassow [BKT12]. DASWIS [Ari02]. Data [ABB12, ARSB09, AGL^{+}07, ASR^{+}07, Atl08, AHSC11, BdBC01, BA07, BP11, BLB^{+}03, BSTL07, BGS^{+}03, BDC12, BH00, CGL^{+}09, CBG^{+}10, CFZ10, CZF10, CBS^{+}12, CHB^{+}12, CUDL02, CWiM^{+}09, Che99a, CLY02, CK06c, CSN14, CAB10, CGS02, CK07, CYBF06, CY09, DSZ04, DL06a, DDS12, Dan12, DST01, DSS03, DSZ07, DGST09, DIS11, DR06, DD07, DGDM^{+}11, DG09b, DFS08, DFM11, DFT12, DLW^{+}07, DFGN12, EMT02b, FKM08, FKLY08, FTC^{+}10, GVO^{+}10, GBH11, GVB12, GG98, GZSC07, GJS08, HKW09, HKW^{+}10, HKW11a, HKW^{+}11b, HKW12a, HKW^{+}12b, HCL05, HKT12, HYP^{+}09, HCS11a, HCS11b, HdMKS12, JW05, JP05b, JP06, JP07b, JP08, JP09b, JP10, JP11b, JP12a, Jür02, Kam99, KMW03, Ker12, KMSS02, KT10, KZ10a, KŚHM11, LwCM00, LZhL06, LFP^{+}09, LM03, LCY03, Mac12, MCKB06, MM08, MMM10]. Data [MLK04, MF06, NKLC06, OMMSP06, Per02a, PBH03, PI05, Per07a, PS07a, Per07b, Per08, PS08a, Per09a, Per09b, Per10a, Per11a, Per11b, Per12a, Per12b, PRG09, PRG10, PRG11, PdA^{+}11, Qui04, RTZ08, Rei99, RtTMP10, SST00, SPG10, Sad13, SCH^{+}09a, ST08, ST11, Shi06, Shi07, Shi09, Shi10, SEK^{+}11, Shi12, SK12, SBM08, Sio10, SCD^{+}10, SWS^{+}05, SYP^{+}05, SMA03, Spa03, SBJ^{+}05, SAC^{+}06, SACM06, Spa06, SAF^{+}07, Spa08, SPT^{+}08, Spa09, SZS09, SD09, Spa11, SB12b, TCHB12b, TCHB12a, TLZ^{+}08, TKCW11a, TKCW11b, TKCH09, TNA^{+}10, Tsu08, WZH^{+}07, WSTI08, WM04, WPZA11, YEY04, YAF^{+}02, YTC^{+}07, YWRS11, YCB12, ZSD03, ZSSM03, ZH00, ZYRP10a, ZYRP10b, ZZZL12, ZLY07b, ZKZ00, Akl95, BMSMS04, Bha95, BC93, Bre03, Che03b, CWL01, CMM01, Dam99a, DF01a, DSS89, DSS91, Deh93, Deh97]. data [Deh99, DLOS05, Ehr91, EO94, FKP^{+}05, GHM05, GJ86, GLS05, GF12b, GSV05, Gun88b, HKB99, HOD96, Hil78, Hof01, JP04b, KMP84, KS95, KZ97, LG94, Leo00, LWD05, LS89, LCB97, Lom93, MMR07, Mul93, Ove83, PZ92, PY05a, PP99, Per01, PR03b, Per05, Per06, PD96, RTA95, RH01, ST88, Sch97b, She85, TLF01, TLC00, US90, Wan03, Wha03, WGL96, WKK98, ZZ99, ZSO99, ZS05a, ZQ98, ŻR99, dRS01, vM94, Atl08, BCBF08, BA07, Bla80, BR12b, CBT07, CK07, CBCGA12, CD11, CD12, DKY06, DL06a, FJ10, GANACBR10, GANACL11, GANACBdCdV12, GV09, Ham08, HT09, HMT10a, HT11, HHMT12, HL05, HS76, JW05, KMT00, KWA01, KWA02, KMW04, KA12, LK10, LSgL^{+}06, LNE06, LYT06, Li11b, LR05b, Mac12, MT99b, MO92a, PMH09, PMT09]. Data [PMT10, Per02b, Pie05, Qui04, Rah04, SXC04, SEN07, SEN08, TT05, TT06a, WS06, X^{+}12a, Yao82, YK82]. Data- [HKW09, HKW^{+}10, HKW11a, HKW^{+}11b, HKW12a, HKW^{+}12b]. Data-oriented [KT10]. Database [AK90, AB12, AYBH^{+}06, AH98, Aus86, BBB^{+}07a, Bel03, BMR^{+}04, BHRU09, BCP^{+}04, Bis88, BCH^{+}05, CM00, CFB11, DT89, Gae96, GL10b, GHI^{+}06, GPV88, HKW^{+}11b, HKW^{+}12b, ISS^{+}06, KKMN07, LYBU10, LM08b, NLŚ10a, NLŚ10b, NKJ11a, NKJ11b, PCN12a, PCN12b, PCN12c, ŠPTM00, TDG91, Win09, AB93, Ame91b, ABG96, BC89, BTT90, BJS99, Cas81, Cha03a, Dit88, FNY93, GB97, GL92, JBJ94, JJP^{+}02, Kam99, KS95, Lev92, LMT^{+}04, Mic90, PDG92, RKS89, SST00, SSRA98, SCM88, Sch91, Sch97b, Sel95, Zan00, AK97, AYBH^{+}06, ADW05, AS07b, BdBC01, BBB^{+}07a, BB99, Bel03, BMR^{+}04, BHRU09, BCST99, BKW08, BKW09a, BK05b, BH92, Bla90, BF89, BCH^{+}05, BKW06, BHQ10a, BHQ10b, CLM03, CS87, CLLD09, CH98, EL04, GTT04, GG09, GG02a, GV95, HT97, HCT02]. Database [HLSZ11a, HLSZ11b, HKP08, IKR00, JNS05a, Kar94, KILW10a, KILW10b, LS04a, LG94, LLWL04, LTW06, LYBU10, gLPZ^{+}12a, gLPZ^{+}12b, LSTZ12a, LSTZ12b, MLW93, MRŠ03, May01, MLK04, QSBC98, RT95, SS06a, WT96, WRP07, WGL96, XYZU11, YMY^{+}10, YKU11a, YKU11b, YYH^{+}12, ZOM05, ZYDL09, vdBV01]. Database- [HKW^{+}12b]. database-related [Cha03a]. Databases [ACJ08, BBGS10a, BBGS10b, BBGS10c, BH06c, BDG04, BB06, BEGP04, BRM06, BGKS04, BGMST09a, BGMST09b, CIT11, CCG08, Col07, CK07, DGM08a, DGM08b, DZ10a, DF02, DFT04, DFF06, DFS08, DFM11, DFT12, DS07, ENP02, EHKP05, EBT11, FGB11, FDC12a, FDC12b, FSS06a, GL92, GJS08, GMV09b, GKMN10, GHMV11a, GHMV11b, GHMV11c, Had03, INR07, JNS05a, JLY03, JTB^{+}05, KMTW03, KKdM^{+}07a, KKMN07, KSF^{+}03, KR04b, LGTB03, Mac12, MSP^{+}09, MN02a, MPS06, MEB05, MHW12, NG10, PZK07, PTM^{+}11, Sex09, WM04, AS99, Aso04, Atz93, AMM99, Ber03, Bos96, Bow94, BRR97, Buc90, BTF05, CE01, CTT93, DKM91, Dit01, ERW99, EH95, Emb98, EJS98, Fre98, Gae96, GK95, GS05, GW93, GS91, GPL99, Jen01, KW96, LMV95, LJ00, LR94]. databases [LMV98, MK96, OA93, PZ96, Rea01, SV97, Sma97, ŠPTM00, SV01, Tuc01, AKK04, Bha00, Bha02, Bha05, Bha07, BB03, Jon00, Jon01, KSB10, KMSB11, MBJ^{+}07, NOOS05, AFS89, Gra91, TL98]. Datalog [BP12a, dMGFS11, BP12a, dMGFS11]. Dataspace [Sex09]. Datatype [BGHJ07]. Datatype-Generic [BGHJ07]. DATAVIEW [MDH10c]. DataX [GHI^{+}06, LMT^{+}04]. David [Gol11b]. Davis [ET01b, IL01, LMW01, SW04]. Davos [Gün88a]. DAWAK [PMT10, CD11, CD12, PMT09, SEN07, SEN08, TT05, TT06a, KMT00, KMW04, KWA01, KWA02, KMW03]. DaWaK’99 [MT99b]. Days [Reu01, Reu97, Reu99]. Dayton [BG11a]. DB [LMT^{+}04]. DBIR [CLZ^{+}09]. DBIR-ENQOIR [CLZ^{+}09]. DBISP2P [MBJ^{+}07, NOOS05, AKK04].

DBMAN [CWC^{+}07]. DBPL [AS07b, GG09, BK05b, CH98, GG02a, LS04a]. DBPL-6 [CH98]. DBPL’99 [CM00]. DBSec [CBCGA12, Li11b]. DC [Can12, DHRT10, Fei03, GP09b, HdMKS12, PS07b, SW03a, SF08c, AKZ01, BS01c, FT04b, JS01]. DCE [Sch93a]. DCFS [HKP11, KMR12]. DCOSS [ASAM07, GAA06, KSHM09, NCJK08, PISW05, RMDS10]. DCW [Kro00, PKSS02]. DDEP [KM04]. Deauville [BF89, MS12c]. debugging [Fri93, DHM00]. Dec [JS84]. Decade [BL04b, dBdRR94]. December [ARS03, ARS09, Abe10, AK90, AR99, AS02a, AF97, AdBB12, A^{+}08, ANS04a, ANS04b, ANS04c, AK03, ABdJS10, AKR^{+}12, Alo05, APBP07, ACK00, And94, APW06, ALR11, AM98, AF99b, AM00, AJS00, AKG06, AR98, AP07, Asa96, Asa06, AiNOW11, AKSO06, BC09a, BPSP06, BA06, BBT08, BKS99, BLO^{+}07, BYLJ11, BSW95b, BL06a, BS11d, BS11e, Bau99, BBKP06, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BCT06, BCG^{+}07, BV92, BCD05, BC11a, BRL08, Bha00, Bha02, BN91b, BTBP^{+}10a, BTBP^{+}10b, BC07, BCSW10, BNS00, BAL88, BP04b, BKM08a, Boy95, Boy01, BL07b, BJ01c, BC09b, BJS93, BRR97, BHM03, BMC08, BHH08, BOH^{+}12, BB00b, CLMP02, CCD04, CV04b, CYGL04, CSJD06, COT98, CT06, CFM09, CPSA10, CSW06, CD07, CTT93, Cer07, CW11, Cha05a, CV96]. December [CL06a, CDPT06, CMM^{+}09a, CCH02, CCM^{+}04, CY10, CTS10, Che11, CEK11, CY11a, CFLR12, CKLN10, CCP10a, CCP10b, Chi95, CBS^{+}11, CW94b, CJ05, CRD08, CKM^{+}09, CI98, CFVV^{+}12, Cod01, DYCD09, DL06b, DD97, Dar97, DB02, DD03, DSDS04, DG04, Dat09, DBC^{+}10, DGG^{+}11a, DKM91, DQBZ02, DD06, DY06b, DG07b, DsSL^{+}07, DWML05, DY10, DJS91, DFF06, DA08, DDI09, DvV96, DW07b, DHH^{+}08, ET01a, ELS12, Ehr87, EKT94, EYX^{+}05, EIRW06, ER12, FSA^{+}07, FV04, FV05, FLY06, FL09, FPMW06, FT04a, Foo99, FNPW05, FHW08, FW92, FLMS05, FLT07, FKLT08, Gal07, GN12a, GMO09, GF03, GLP99, GRB^{+}07, GKT09, GKT10, GKST11, GKS12, GM04a, GDP07, GPM98, GL74a, GL74b, GMI06, GCSR07, GPYD09, GG10, GSV05, GYD^{+}06, GM10, GDG11, H^{+}05, HMV01, HP95, HM12b, Heb95, HWG10, Hig05, HJW00]. December [HZ08, Hon01, HNF08, HNS^{+}12, HO98, HL91, Hu09, HXC^{+}08, HL00, HS08c, HMCL06, Iba92, IKO03, IB07, IAL^{+}07, IGN05, Ish97, IS08, IZP09, JZR09, JY96, JM05, JM11a, JRL09, JM07, JM02, JLS11a, JLS11b, Jha10, JM11b, JWH06, JRJ05, JM03, Joh97, JS11b, JMO10, KP06, KL95, KCYL08, KUA00, KP00, KMMN11, Kap08, KR02a, KR04a, KMSB11, Kim02, hKhLKŚ10, hKAS^{+}11, Kim12, KŽ04, KK11e, KRH09, KS10b, KSG^{+}07, Kur07, Kur10, LC06, Lai03, LT00, Lee04, hLhKcFŚ09, LC09, LH10, LW11, Leo09, LIJ97, Leo00, LwCM00, LLC09a, LLC09b, LPC^{+}11, Li96, LLT^{+}04, LSDN04a, LSDN04b, LV05, LKZ^{+}08, Li11a, LSS^{+}04, LFK^{+}02, LTW11, LXH^{+}12, LMV95, LYL06, LYhL^{+}01, dVLKR^{+}08, LM04, Lou94, LMM10, LSW^{+}10, LCY^{+}11, Mad89, MCK^{+}06, MCKB06, MWYF10, MT09b, Mah04, Mah85, MPP08, MMV06, MBW08, MSY08, MVCM10]. December [MLM90, Mat09, MSWZ06, MTO99, MRY^{+}11, MSW05, MG07a, MS02a, MS02b, MR07a, M^{+}06, MPV01, Moo01, MWK^{+}09, Nak92, NcC99, NSM07, Ng93, NL09, NV01b, NQL06, Nor86, Nor87, NK88, Nor90, OKS02, OS07, Oka00, OBS12, OT07, OWPY03, OALZ^{+}08, PR01, PWZ98, PBB05, PKKG06, Pal03, PJC^{+}12, PCH^{+}06, PCK^{+}12, PDSD10, PSDS11a, PSDS11b, PDSN12, PZ08, PH04, Pap95, PA09, PC05, Par09, Pat03, PTB99, Pic01, PSN95, POS00, Pie08, PP03, PA03b, PMC06, PG09, Pre95, QMLW05, QIW07, QMW09, Ram08, RRR99, RS10, RCA90, RS06, RS97a, RAW07, RD01, RSS06, RL06, RN11, RK09, RPBP06, RSvdP05, RO00b, Roy05, RS09c, Sab10, SPBP08, SKS02, SCH^{+}09a, SSO^{+}11, SÇS12, Sar03, Sat97, ShK06, SH91, SZ93, SP08, SZC^{+}03, SC09a, Sha07, SKK08]. December [Shv06, Shy92, Shy93, SU97, STA^{+}11, Sko85, Sma05, SAUL11, Son00, SQL10, SMK06, SH09, SRY07, SB12b, SKCK07, Sta95, Ste01, STTP02, SCB01, Sub05, SLCT11, SV05, TWYhL01, TKCW11a, TKCW11b, Thi94, Thi95, TPCB12, Tok07, TNI09, TTF07, Ued10, VS09, VKV00, VAR12, VG12, WSS04, Wal99a, Wal01, WR11, WS12, WY99, WY04, Wei97, WHG07, WZ08, Won01, WK06, WCGB05, WD10a, WD10b, XLKZ12, XXX^{+}09, XL11, YZZ^{+}05, YAL^{+}05, YYI^{+}09, Yan11b, Yan11a, YT99, YTC^{+}07, YLL09, ZTTY12, ZCRL08, ZHF04, ZJ05, ZOCJ07, Zhe02, ZLY^{+}07a, ZF05, dlBP08, vSH06]. December1 [RK06]. Decentralised [ABD^{+}12]. Decision [BLvdH10, BRT11, CMSM^{+}10, HSL^{+}99, Mah04, MGSM12, RT09, SS76, THL11, TNYL11, Bar03, BHR84, Bui87, Fer93, ABB10b, BKA11, GW12]. Decisions [TN04, TNM05, TNVDF06, TNY07, TN08, TNI09, TND10]. declarations [SS89]. Declarative [AS11a, CP10b, GS09b, Han07, Hen06, HW08b, Jay04, KR02c, PC00, RL11b, RZ12, SHW09, UWB^{+}06, DW03a, GHK97, Gup99, HC05, Nad99, PGM98, Ram01, RB91, SHGB05, BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, LOST04, LOTY05, OSV11, SSVW12, UWB^{+}06]. decoding [Gur04]. Decomposition [Hel08]. Decompositions [Kai11]. Dedicated [AG09, ADM11, AAN09, ADR08, BS11b, BDR10, BDFM12, BKT12, CRS11, CLKK07, CS12b, DDM08, DKN12, DKS12, EMO10, ELS^{+}10, GKKW09, HL05, Hei12, JL11, JPR04, KK11c, KSW03, KR11, Len04, LLM09, LIBP07, Mah04, MM09a, MRU12, MSS02, Nac12, Pal09, PQR11, RSB07, SS06e, BEKS02, FJM06, HNR05, JRR03, KMPR04, KMO^{+}05, MvOvRdV05]. Deduction [Baa03, BK80, BSS11, BR07a, CS00a, Lov82, LO88, Pfe07, Sch09b, SNRG11, Sho84, Sie86, Sti90, SZ11, Vor02, Win04, Bun94, Gan99, GWY99, HW06b, Kap92, KK98, McA00, MP96, McC97, MS96b, Nie05, RGW01, Wan98, Wil91]. Deductive [Fre98, MR12b, Atz93, Sch91, BRR97, CTT93, DKM91, LMV95]. DEECS [LSgL^{+}06]. Deep [OFK05]. Deerfield [AN09]. Defence [TGH06]. Defense [LSW12]. Defenses [BSC12]. Definite [HR99b]. definition [FNY93, BJ84]. definitions [DJL88, Don76]. Deformable [PH02, PD04, PF06, PF08, PF10, PFM12, NL01]. Degree [Sta99]. Del [GYA^{+}06, Bar08a, BVAT12, DZGS06, EL07, FdS12, GPV00, Ish97, KSHM09, MÁ07a, MÁ07b, NVSB^{+}07, PISW05, WB09]. Delage [CFG96]. Delaunay [GTM11]. Delft [ABR09, BMSHW09, BCV06, BTPPS07, JSWhT11, KHPS07, KH12, LV07a, LAF^{+}10, LWRdR11, MvL10, NdG04, SEL09, TMdB11]. Delgada [CCD^{+}09, FLMG10, MZF94]. Delhi [AP07, BS11c, BN91b, CMM^{+}09a, Gel12a, Gel12b, HKP08, Jha10, JM03, KP00, KK11b, Mah85, MG07a, Nor86, PA09, RM04b, RS09c, SCH^{+}09a, Shy92, SB12b]. deliberation [HWP01]. Deliberative [Har11b]. delivered [Der04]. DELOS [TBC07, TAS05]. Delphi [AK97, Gae96, TSK92]. Demand [BCL03, LY01, SF10]. DEMM-2000 [BdBC01]. Democracy [BGPW05, KLLP12]. Demonstrations [FMC12a, FMC12b, FMC12c]. Demos [DB98b]. Denmark [ADW05, ACL98, AMCL02, AYB^{+}09, AH97a, Arn90, BG10a, BC95, BJ10, BG12a, Bri95, BL02b, BFMS01, CF09a, CDCJL08, CSV05, Cra05, Dam91, DF01a, DW94, DMW96, DMW06, EF02, EL02a, EP07, FT11, FS09a, GJ86, GM01, HK04, HMT99, HS02, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, HSL^{+}99, Hut96, JR98, Jon80, Jon90, Jou11, Kin08, KRRBV07, KMYJ08, Kuc11, LS92, LSW98, LN01a, LNS06a, LNS06b, LPK05, MSP^{+}09, MPB05, MFG12, Mey01, MNS95, NS82, NT98, NS00, NZ03, NF07, NSFB10, Nür03, Olt95, OALZ^{+}08, PM96, PDF^{+}09, PHOK10, PBW05, RR98a, RR01, RTB02, SHØ06, SS94, SJDT08, Sta82, ST12c, Stu02, Tis02, TBJ02, TT05, Tou94, VTK^{+}12, Vid12, Vor02, Waś96, Wii05a, WTGA05, WL93, YKMJ09, YAKMJ10, ZBH12, dGN12]. denotational [Bli87]. Denver [Gho10, Lee01, SS09b, vdBGG10]. DEON [ÅBE12, GM06a, GS10, vdMvdT08, LN04b]. Deontic [GM06a, LN04b, ÅBE12, GM06a, GS10, LN04b, vdMvdT08]. Dependability [MCER10, Sch12, EFR08]. Dependable [BTF02, CdLG10, EHP94, GNR07, HMP99, HSS96, Pfl02, dLGR05, dLGR07, dLDG^{+}08, dLFG^{+}09, DKP05, dLGR03, dLGR04, BTF02, BBR07, DKP05, EHP94, HMP99, HSS96, JL11, KMS06, MSAdAS05, PvS07, SNČ^{+}12a, dLWC03]. Dependency [Kuh10]. Deploying [BHMS07]. Deployment [Bis02, DKK08, EW04, HTW02, DE05]. depth [dB93, NMS93]. Derby [APHA11]. derivation [Mul93]. described [O’D77]. Description [Fen00, Küs01, BO80, BLW82, Bus85, Bus88, Haj96, KBD87]. Descriptional [HKP11, KMR12]. DESEC4LCCI [OD12b]. Design [AO02, AGM05, AG07, ABG09b, ABDW10, AG11, AJB09, AGCP^{+}11, Ayk09, AS07c, BR12a, BI03, BH07, BDZ08, BPZ09, BI11, BCP12, BAdVR08, CBF09, CKK^{+}09, CM03b, CB06, CLS07a, CKOB11, DZGS06, DB07, EPVW10, FS07a, FD03, Fed01, FMS02, FG01, FG04a, FLUV02, GNR07, GT03, GH06, GW98, GP08b, GI03, HP02a, HOC07, HAB02b, HKH10, HJ00, Jac07c, Jac11a, JSV11, JJF03, Kur09, Kur11, LK99, LWZ09, LYCV08, Luo04, Luo06, Luo07, Luo08, Luo09, Luo10, Luo11, Luo12, MRC12, MPK04, MSB12, MB06, MvL10, NM08, NSFB10, OB07, PBCF08, PB08b, PRR05, RCC12, Rau11, RBG09, SH12b, Sch01b, SCL^{+}06, SLL^{+}07, SYY^{+}08, SS07c, SPB00, SO99, Ste11e, SKM07, SM09, TNP^{+}07, VAM06, WJP07, WZA10, dVGR04, vLS11, BC89, Bee89, BC04a, BB05b, BP05, BS05]. design [CE95, CYR02, DTV02, Enc80, FKS86, FH01, FHM82, FLS05, HM04, Joh01, KDS01, KN83, KK95, KBD87, Lam96, Luo05, MSRP98, MM01, Mar80, MP93, MT99c, Mul96, Ove83, PP01, PVV05, Pau93, RR03, RMV96, SLB05, Smi98, ST87, SC96, Ste05, Tob80, Van85, Yao82, YK82, Buc90, CL96, DBSvD09, Dig11, ER12, Fit93, Mar11a, Mar11b, Mio90, Mio93, PRK12, PS93, Rei03, WF77, ZS00, Pet80]. Designing [FPRAW04, LSJ06, SS09c, SBH07, SBH09, BB96, PSP^{+}02, AN06, Fed01]. Designs [PM08, RB08]. DESRIST [JSV11, PRK12, WZA10]. Detection [DMW00, Fri06, HAB02a, JSK10, JVA04, KJB09, KLC07, LKT08, MBS05b, Pfl02, SBM11, VKJ03, WVD02, ZK06, Don89, JK05a, LMW01, VZ06, BL06d, FB09, HS07, HB11, KJ10, Zam08]. deterministic [LT01, WNL05]. Detroit [KS08b]. Deva [WSL93]. Developing [Ber99, KRMM^{+}02]. Development [AM91b, AM91a, AW12, Ayk09, Ber02a, BM10, BJ78, BBPP09, BDGK00, BKM08b, BL04b, BJRT06, CJ11, CG12, CR01a, CKK^{+}09, CS10, CM09c, CDFP11, DSLT08, DSLT10, DSHL11, DO89a, DO89b, Ehr85a, EEK^{+}87, EKLM87, Fen00, FS07c, GDG11, HW07, Jac11a, JK06, KOFJ09, KMK10, KMSJ11, LCH^{+}12, MOP11, MOH^{+}12, MW09, MJ07, MK12, MAR05b, PVW10, PRdRW09, PT91a, PT91b, RW11, RA06b, RA07a, RA07b, RO09, Rau11, TBC07, WPR08, WGMP09, WR07, Wes99, WWD99a, WWD99b, WPH03, ZYBX08, AW01, ABB99, AB03, ABGP05, Ber98, BCM00, Ber02b, BBKW03, BR97, BKL^{+}91, BD97b, BHL90, Ehr85b, EW91, Fia99, FMO05, GJ93, HM87, HKB93, LP97, LL95, MPS93, MNS95, MD01, Nag96, Par98, RA06a, SS83, SLF91, TKHS05, VA95, WSL93, WKK98, ZTY^{+}05, Zho01, ZLS^{+}06, BL04b]. Development [CW07, ECV^{+}10, vdL98, BS05]. Developments [BHVV09, BBK^{+}11, DMB^{+}11, Li96, Wet92, CCD04, CB11, CHA03b, DR05, DN09, EF03, GLSY10, HKL07, HA02, ID06, IT03, IT08, KRS02, ML11, WV99, YI12, vL95]. Devices [AZ11, BGH06a, Chi03, Fer08, IS12a, IS12b, MBH^{+}09, OSC^{+}08, STP^{+}10, BBH^{+}05, Pat02]. DeWeb [IHX^{+}08]. DEXA [ADW05, BKW08, BKW09a, BKW06, BHQ10a, BHQ10b, HLSZ11a, HLSZ11b, LSTZ12a, LSTZ12b, WRP07, GTT04, HCT02, IKR00, Kar94, MRŠ03, May01, QSBC98, RT95, WT96]. DEXA’97 [HT97]. DEXA’99 [BCST99]. Dexter [CS12b]. DFG [EDD04, MO92a]. DFT [Sch12]. DFVLR [DMWW88]. DGCI [ADL05, BRP09, CSTD08, DRDKE11, KNP06, AF97, BCP99, BNS00, BLV02, MMU96, NSS03]. DGOV [WTZ^{+}04]. Dhaka [iNR08, RF10, RiN12]. Diablerets [Vau05]. Diagnosis [BKNS06, MDY^{+}10, Nok91]. Diagrammatic [BPCS06, BMS04b, CPR12, GJN10, HNM02, SHL08]. Diagrams [ACH00, BPCS06, BMS04b, CLS07a, CPR12, EPVW10, FP08, GTA10, GTM11, GJN10, HNM02, SHL08, WJP07, Kle89]. dialogs [DEM81]. Dialogue [ADMH04, ADM^{+}08, CPSA10, HM11a, LT11a, LMMN10, MM03, MM07, MM09b, SKP00, SKP02, SKP04, SKP06, SHKP08, SHKP10, SHKP12, MMOS99, Mat01, MMP05, MML97]. Dialogues [CSCR08]. Diamante [BGLT05]. DIANA [Goo83]. DICS [KMS06]. Dicsovery [AS02b]. dictionaries [GP89]. Diego [Ayk09, BC07, BKM06, BKM08b, BG07b, CS05a, Cha02b, Cou03, DG07b, Duf09, EW05, EIRW06, FV04, Har09, IB07, Jac09c, Jac09d, Jac09a, Jac09b, Kar09, Kur09, Lin14, LR05b, MS00a, MZC06, OZ09, Por98, SS09a, SEG09, Shu09, Sle92, SS09c, Ste09b, Ste09d, Ste09c, TD07, VH04]. DIEW [YMY^{+}10]. Difference [dRLP98]. differences [Tha81]. differential [Chi79]. differentiation [Ral81]. Digital [AT02a, ABK^{+}09, AKSO06, ALL^{+}10, BCG^{+}11, BB00a, BC02, BMC08, CCM^{+}04, CC12, CKH10, CDCJL08, CKM^{+}09, DGG^{+}11a, Duf07, DHH^{+}08, FSA^{+}07, FHFL06, FHLP09, FHKQ12, FNPW05, FKL08, FLP11, GSH^{+}04, GMI06, GCSR07, GTVC06, GBMS11, Gun94, cHPkC^{+}07, IZP09, KLP04, KLS10, KM04, KS04b, KFM07, LJR^{+}10, LPT07, LFK^{+}02, LPK05, NS98, Pal01, PAD^{+}06, PZE^{+}08, Pla96, SZC^{+}03, SBC^{+}94, STA^{+}11, SS08b, SHRM06, SBH^{+}04, TvdBI02, TSE^{+}09, XCR11, ZBRL12, ZZP^{+}10, vdBK05, AV99, AYB95, AS99, CS01a, HL04, Her92b, HAB98, Klo87, LG95, MNW98, PT97, Ran93, RCT05, San02, TAS05, AYB95, Ada96, ABRZ06, BCKK05, BBEdK98, BBGR03, BIK01, BK81, BMC08, Cai08, CLS11, CCM^{+}04, CKL05, Duf07, Duf09, Duf11, Fei03, FNPW05, HSKB09, IFGH10, II00, KCR04, KKH09, KSB11, KM04, Kru08]. Digital [LFK^{+}02, MOFM10, PK03, Pfi96, SNY06, SZC^{+}03, SJ06, SKK08, SKPG12, SHRM06, TvdBI02, TBC07, Wes01, XCR11, vdBK05]. Dijon [CDV03, Jou85b]. DIKU [HMT99]. DILS [BCBF08, BR12b, CBT07, LK10, LNE06, LR05b, PMH09, Rah04]. DIMACS [TT95, IWC04]. DIMACS/RECOMB [IWC04]. Dimensional [Yu02, Yu02]. DIMVA [BL06d, FB09, HS07, HB11, JK05a, KJ10, Zam08]. Dines [JLW07a]. Direct [OZ83]. Directed [Fri06, Jon80]. Directing [SF11]. Directions [ABF^{+}07, BRS^{+}07, CJR^{+}10, DMB^{+}11, DsSL^{+}07, LS12a, MG07b, ZSY^{+}06, BL92, BFS92, Pla96, Pup99, Rob93, RBS93, Sch03a, ZSO99, Bow94, CB10, HLR00, LSW81, Che99b]. Disabilities [AGB^{+}11]. Disappearing [SKM07]. DISC [Agu12, Dol06, Fra05b, Gue04, Kei09, LS10b, Pel07, Pel11, Tau08, Fic03, Her00, Kut98, Mal02, Wel01]. DISC’99 [Jay99]. disciplinary [SAUL11]. Disciplines [CDS09]. DISCO [CL96, Fit93, Mio90, Mio93]. Disconnected [Kis95]. Discourse [Bra07, DBM09, HDBM11]. Discovering [APHA11, BHMS07, MLK04, ZSSM03]. Discovery [BBGS10a, BBGS10b, BBGS10c, Ber12, BB06, BEGP04, BHH^{+}06, BGMST09a, BGMST09b, Cai05, CLY02, CD11, CD12, DGM08a, DGM08b, DSZ04, DT07, DS07, EMT02b, FDC12a, FDC12b, FSS06a, GVO^{+}10, GHMV11a, GHMV11b, GHMV11c, HAB02a, HCL05, HCS11a, HCS11b, JTB^{+}05, KMT00, KMW03, KMW04, KKdM^{+}07a, Lac10, LV12, LSS02, LGTB03, MS10d, MNLM06, NZS^{+}06, NSS^{+}07, NZ06, NKLC06, PMT09, PMT10, SEN07, SEN08, TCHB12b, TCHB12a, TKCH09, TT06a, TWSN07, WJ05, WJS^{+}06, WZH^{+}07, WSTI08, ZYRP10a, ZYRP10b, ZSM^{+}09, ZLY07b, ZKZ00, AS02b, CWL01, GS05, KWA01, KWA02, KW98, KZ97, MT99b, TLC00, TT05, Wha03, WKK98, ZZ99, ZQ98, ŻR99, dRS01, AM98, AF99b, AM00, CTS07, EHM11, GCJB09, GLP12, GTY03, JS01, JFBH08, MSH05, PHH10, SA04, TLJ06]. Discrete [Alt01, Dad01, KMS12, Tha84, AF97, AKU01, AK03, AKT06, ACK^{+}07, ADL05, BCP99, BNS00, BLV02, BRP09, CDV03, CSTD08, DRDKE11, KUA00, KNP06, MMU96, NSS03]. Discussions [MW08]. Disease [DPNM12]. disk [Ful75]. Display [YAKMJ10]. Dissemination [XCR11]. Dissertation [Lib07, Gur04]. Distance [Ahn04, Ruc96]. Distributed [ARS09, ACK00, APW06, ALR11, BBT08, BdB08, BG09, BR97, BMO02, BJ07, BD11, BK04b, BSW12, CCS04, CGS07, DG06b, DKK08, DPV11, DV07a, DCFY10, DMV01, FKB02, GR12, GW06a, GAR04, GRR05, Gue01, GP09c, GYD^{+}06, HZ10, Hig05, KHW03, KS00, KS08b, LLPH09, LMM10, MT07a, NM08, NNS03, NPPDG06, PWZ98, PH04, PV02, RUL04, RRU05, RS12, SW04, Sar03, SvS00, SS05b, Shv06, SS99, SvdMOP06, STY^{+}07, SHYEF08, SC00, TTF07, Wan05a, WPR08, ZL97, ZL98, dFEN04, AM91a, AW05, BGW93, BMS95, BGY96, BMP96, CYGL04, CPGL05, CFP94, DOS99, ET01b, FRA05a, GP98, GAR03, GNR94, GY03, Ish94, Kis95, LM98a, LSS^{+}04, LPH00, Neh88, Nok91, Pad99, PCG^{+}05a, Rol98, Rol99, Rol00, Sch93a, Sch03a, SPZ90, She01, SS90, SLL01, SLPM96, SZ05]. distributed [SW97, VJ99, X^{+}12b, dBdRR90, von79, AYV^{+}11, Agu12, ASP85, ASAM07, BM96, BR89, Bod91, BDDM12, CCS04, Cha05a, CDPT06, DB02, DD03, DSDS04, Dol06, DSS^{+}07, EM06, EK10, FR11, Fic03, Fra05b, GWK09, GM04a, GAA06, GH12b, Gre94, Gue04, HR95, Her00, HGZ05, IR07, JM07, JM10, Jay99, KPSW10, Kei09, Kra97, KSHM09, Kro00, Kut98, KA05, LSDP81, LPSD83, LS10b, MCK^{+}06, Mal02, MT97, MT08c, NO11, NCJK08, ORB^{+}09, PKKG06, PA09, Pel07, Pel11, PM96, PKSS02, PISW05, RMDS10, RR12, RCJ^{+}08, Sch93b, SZ92, SO09, SDH03, Tau08, TV94, TSK92, Wei97, Wel01, ZL96, vL88, vLS91]. Distribution [CvSW04, CNY95]. District [HK07]. Diversity [Ste07d, Ste09b, Ste11f, Ste11d, MD01]. DIWeb [Bre03]. DL [Har11b, AYB95]. DLT [CCD04, DR05, DN09, GLSY10, HKL07, ID06, IT08, ML11, YI12, EF03, IT03, KRS02]. DMG [Pie05]. DML [JCC^{+}12]. DMTCS [CDV03]. DNA [CP06a, CS11, DS09b, FMZ05, GY08, GSS09, MY06, SM11a, ST12c, CP06a, CS11, CR04a, CR01b, DS09b, FMZ05, GY08, GSS09, HO03, JS02, MY06, SM11a, ST12c]. DNA-Based [CR01b, JS02]. DNA10 [FMZ05]. DNA11 [CP06a]. DNA12 [MY06]. DNA13 [GY08]. DNA8 [HO03]. DNA9 [CR04a]. DNIS [Bha05, Bha07, KSB10, KMSB11, Bha00, Bha02, BB03]. do [CDGW10, CEM^{+}06, dRCVT11]. DOA [MT05a, MT06a, MT06b, MT07b, MT07c, MT08a, MT08b, MDH09a, MDH09b, MDH10a, MDH10b, MDH^{+}11b, MDH^{+}11c, MTP^{+}02, MTS^{+}03, MT04c, MT04d]. DOA-SVI [MDH^{+}11b, MDH^{+}11c]. DOCM [DHtM^{+}12]. Doctoral [GKMN10, HK12, KBČ^{+}12, Lib07, MTH06a, MTH06b, MTH07a, MTH07b, RKM^{+}12, Gur04]. Document [IS12c, KM04, Uni81, MNW98, MB97, NW97, BS06c, wLN99, LHK02, MD04]. Documentation [Mos04, Des86, ZWS81]. Documents [FKLT08, KM04, Koh06b, NZ06, Vaz99]. DoD [WF77]. Doha [BJM09, Cer07, CVV08, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e]. d’Oleron [GP89]. Dolny [Tar91]. Domain [EK09, Hel08, QBY^{+}12, TQYX11, Hal96, Wal99b, BB07, GLW07, LBCO04, Tah09]. domain-independent [Wal99b]. Domain-Specific [Tah09, LBCO04]. Domains [Jac09b, JP03, WTZ^{+}04, Rei99]. Dominica [PY05a]. DOOD [BRR97, CTT93, DKM91, LMV95]. Dornburg [BKM07]. Dortmund [BCD^{+}03, BBAB^{+}07, CK82, HKH10, Müh75, Reu97, Reu99, Reu01, RJL^{+}08, SM91, Vah09]. Dosse [Ehr91]. Dourdan [BC93]. Douro [CBF09]. DPA [CAS^{+}09b]. DPM [ED06a, GANACBR10, GANACL11, GANACBdCdV12]. DQDI [YYH^{+}12]. DQIS [XYZU11]. Drawing [BC11c, EG10, GK02, HNQ08, KW07a, Lio04, TP09, Whi98, vKS12, Bra96, Di 97, HN06, Kra99, Mar01, MJL02, Nor97, Pac05, TT95, KW01]. dream [Luk97]. drei [SF74]. Dresden [AJB09, AAB^{+}07, ABW10, Bis88, BL07a, BUHS08, Din03, FR01b, FST09, GH96, HLW08, JHC02, KMTW03, LMSS07, Leh11, MS06b, MdSC09, NWL06, Pfi00, Rei95, WBD08, dMLG03]. dritte [Bra73]. Driven [AVP07, AAR05, BCP12, COB^{+}08, EOSW07, ELS^{+}10, FKBA12, HK05a, KOFJ09, NWHR06, PHR09, PRH10a, PRH10b, PRR05, RW06, SH08, SS09b, WCK11, BW05b, BKM08b, Nau02, PST11]. DRM [Fei03, San02]. DRMTICS [SNY06]. drum [Ful75]. DS [CTS07, EHM11, GCJB09, GLP12, JFBH08, MSH05, PHH10, SA04, TLJ06, AM00, GTY03, JS01, LSS02]. DS’98 [AM98]. DS’99 [AF99b]. DSL [Tah09]. DSOM [BG09, CGS07, DKK08, SS05b, SvdMOP06, ACK00, BK04b, FKB02, SW04]. DSOM’99 [SS99]. DSSCV [OFK05]. DSV [GH06, GP08b, GHP^{+}08, FLUV02, Joh01, JJF03, PP01]. DSV-IS [GH06, GP08b, GHP^{+}08, FLUV02, Joh01, JJF03, PP01]. DSVIS [BPR05, DB07]. Dubai [SSO^{+}11]. Dublin [BMJ88, BGH06b, But10, CFG^{+}11, CF10, DB07, EKO^{+}04, ECV^{+}10, FSN06, FHM82, GGK06a, HKS06, HJMP06, KMF06, LKD08, MH05, OROD07, PMÓ06, Pol01, SC98, SvdMOP06, Ver00a, Ver00b, WAS06]. Dubrovnik [BHM^{+}04, JKN^{+}12, LGTB03, LGBT03]. Duisburg [BBR89]. Dundee [Akm01, Mac12]. Dunedin [PS09c]. Dunkerque [Fon00]. Duration [GLW07]. Durham [BF00, BEFP08, CLMP02, DSLT08, DS06, LG95, MRvM07, QQ11, SPON08]. during [MSY00]. Düsseldorf [LW09, Bau99]. DUXU [Mar11a, Mar11b]. Dyn3D [KK09a]. Dynamic [Han02, Nar01a, Har79, Ove83, Tei93, KK09a, MSY08]. dynamical [And95, VHM07]. DYNAMICS [Fre98, CCBW11, WPR07, ÉEMS04, SST00, Van94].

E-business [LMT00, BHM^{+}02, BFOY04, FKB02]. E-Commerce [LSgL^{+}06, Per08, BPW06, DB09, PW07, PW08, BBP04, FBSMS05, PBW05, BTQ03, FKB02, LY01, NMP^{+}04, Per02a, BTQ02]. E-government [BGPW05]. e-Health [HS11]. e-Inclusion [HM09]. e-Infrastructure [BSW12]. E-Learning [CKK^{+}09, cHPkC^{+}07, PAD^{+}06, PZE^{+}08, CHCM11, ZZP^{+}10, GMUW12]. e-services [BS06d, KMTU03, SDH05, BS03a, BCF^{+}02, CGS01]. E-smart [AJ01b]. E-Voting [AV07, KL12, RS09d]. E. [GG98]. E.I.S.S [BB89, BFS92]. E4MAS [WPM05, WPM06, WPM07]. EA [CML^{+}10, HLC^{+}12, MTC^{+}08, TLC06, Col02, LCF^{+}04]. EACSL [DV10, DH07b, Ési06, GK09, KM08a, Ong05, BM03a, Bra02, Bun96, CS00b, FR99a, Fri01, GGS99, MT04a, NT98, vB97]. EAIA [Fil91]. EAMA [VMvN12]. EAMT [Ste95]. Early [MG07b]. Earth [Wes01]. EASSS [LMŠT01, MŠKL02]. East [ACJ08, BDG04, CE01, CIT11, ERW99, EHKP05, GMV09b, GKMN10, HKM^{+}11, INR07, KMTW03, LPH09, LMV98, MN02a, MPS06, MNW08, MGH09, MHW12, SSZ12, ŠPTM00]. East-European [ŠPTM00]. Eastern [BLVV07, DKN02, MMP03, PVP05]. Ebernburg [Dit88]. EC [BBP04, BPW06, CDS09, DB09, DS08, DKW07, KGG^{+}11, NT06, PBW05, PW07, PW08, RLKHL12, RKS^{+}12c, WKS^{+}10, BMP00, BMP01, BTQ02, BTQ03]. EC-TEL [CDS09, DS08, DKW07, KGG^{+}11, NT06, RLKHL12, WKS^{+}10]. EC-Web [BBP04, BPW06, DB09, PBW05, PW07, PW08, BMP01, BTQ02, BMP00, BTQ03]. ECAI [MML97, NVSB^{+}07, Ria09, RtTMP11, HWP01, MR91, Nar01b, WJ95]. ECAI-90 [MR91]. ECAI-94 [WJ95]. ECAI’96 [Wei97]. ECAL [ARC^{+}07, CFB^{+}05, KKS11a, KKS11b, BCD^{+}03, KS01]. ECAL’99 [FMN99]. ECCAI [LMŠT01, MŠKL02]. ECCAI-ACAI [MŠKL02]. ECCAI-ACAI/EASSS [MŠKL02]. ECCB [BS10]. ECCBR [ABMH08, FG04b, RBGG06, CP02]. ECCV [BS06b, CKMS05, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, FTZ08a, FTZ08b, FTZ08c, FTZ08d, FMC12a, FMC12b, FMC12c, HSL06, Kut12a, Kut12b, LBP06a, LBP06b, LBP06c, LBP06d, PM04a, PM04b, PM04c, PM04d, PZH96, VHM07, BC96b, BC96c, Ekl94a, Ekl94b, Fau90, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, MJ04, San92, SLH04, SKK04, TBJ02, Ver00a, Ver00b]. ECCV’98 [BN98a, BN98b]. ECDL [ABK^{+}09, CDCJL08, GTVC06, HL04, KFM07, LJR^{+}10, RCT05, AT02a, BB00a, CS01a, KS04b, NS98, PT97]. ECDL’99 [AV99]. ECDM [CVLL12, RBsC^{+}06, SPC^{+}08, WTZ^{+}04, OYY03]. ECDM-NoCoDA [CVLL12]. ECI [Sam76]. ECIR [ACR07, BYdVZ^{+}12, BBMSD09, CFG^{+}11, GHK^{+}10, LMR^{+}06, LFL05, MOP^{+}08, MT04b, Seb03]. ECMAST [FM97, HS98d, LG99]. ECMDA [AVP07, HK05a, PHR09, RW06, SH08]. ECMDA-FA [AVP07, HK05a, PHR09, RW06, SH08]. ECMFA [FKBP11, KSGT10, VTK^{+}12]. ECML [BBGS10a, BBGS10b, BBGS10c, BGMST09a, BGMST09b, DGM08a, DGM08b, DGDM^{+}11, FDC12a, FDC12b, FSS06b, GCB^{+}05, GHMV11a, GHMV11b, GHMV11c, KKdM^{+}07b, RTZ08, BdR94, BEG04, Bra93, EMT02a, LW95, LGBT03, LP00, NR98, dRF01, vW97]. ECML-93 [Bra93]. ECML-94 [BdR94]. ECML-95 [LW95]. ECML-97 [vW97]. ECML-98 [NR98]. ECML/PKDD [DGDM^{+}11, RTZ08]. Ecomo [WTZ^{+}04, ALS^{+}05, OYY03, Ari02, JP03].

Econ [HS12a]. Economic [BBGR03, SHS10, Kim01b]. Economics [ANF08, ABR09, CEK11, DG07b, HS12a, Leo09, PZ08, RST09, Sab10, Sch74, SMK06, VA07, DY06b, AR10, CMS11, VAR12]. Economy [LMA09]. ECOOP [Ceb08, Eug09, MØ05a, SC07, Ame91a, Ber00, BHCL87, Bla05, BM98, BBC04, Car03, CNY95, Coi96, D’H10b, Dro09, Ern07, Fro02, GN88, GNR94, HM02, Jul98, Knu01, Mad92, Mag02, MMM01, Mez11, Nie93, Nob12, Ode04, Olt95, Tho06, TWN92, TP94, Vit08]. ECOOP’97 [AM97]. ECOOP’98 [Mag98, DB98b]. ECOOP’99 [Gue99, MD99]. Ecosystems [APP12]. ECOWS [ZJ04b]. ECP [SA97]. ECP’99 [BF00]. ECSA [BG10a, CGB11, MBF08, Oqu07]. ECSQ [KC02]. ECSQARU [God05b, Liu11, Mel07, SC09b, BB01a, CKM93, FK95b, Gab97, NZ03]. ECSQARU-FAPR’97 [Gab97]. ECSQARU’99 [HP99c]. ECSQAU [KS91]. ECUMN [FCLG04]. EDBPM [CW11]. EDBT [GHI^{+}06, ISS^{+}06, LMT^{+}04, ABG96, BTT90, BCP^{+}04, CUDL02, JBJ94, JJP^{+}02, PDG92, SSRA98, SCM88, Zan00]. EDCC [DKP05, BTF02, EHP94, HMP99, HSS96]. EDCC-1 [EHP94]. EDCC-2 [HSS96]. EDCC-3 [HMP99]. EDCC-4 [BTF02]. EDCIS [HTW02]. EDEM [KLLP12]. Edinburgh [AF11, AFS02, ACH00, AFL03, ABM05, ALL^{+}10, BR06, Bod05, BJS99, BJ01b, Bra02, BR96b, Cag00, CG07, Cer05, CFP^{+}09, CSSW04, DS05b, DCP95, EL04, EW04, ER05, FRS02, Fri94, GH10, GW06b, GAN05, HZ05, HMTH10, HdQ08, Jon00, KP10a, KRP06, LSJ06, LOR10, LLLL08, Mil00, MK96, Pol00, RS05, RV11, RBSL06, Sag05, San94, Sas05, SKP10, SS10, TBH03, Tis94, TCJ10, TMP04, WM04, GMW79, PPR87]. editor [Pau93]. EDMCC2 [Bod91]. Edmonds [JRR03]. Edmonton [Ama08b, NW10, SMB03, ZSSM03]. EDO [ET01b]. Education [BBKM11, CKK^{+}09, FKW08, GMUW12, Hol08, IJ06, KM11, Mit06, MS08a, WFZL09, ZZP^{+}10, Dei90, DH94, For88, Gib89, GHK97, HP95, Ibr95, MSZ90, Sle92, Tom91]. Educational [BBDS12, CHCM11, RV98]. Edutainment [CKK^{+}09, CHCM11, GMUW12, cHPkC^{+}07, PAD^{+}06, PCME08, PZE^{+}08, PCME09, PCMC09, PCM^{+}10, PCMY11, PCM11, PCM^{+}12a, PCM^{+}12b, ZZP^{+}10, CHCM11]. EEF [BHK01]. EEF/Euro [BHK01]. EELC [VSTN06]. EEMMAS [WBD08]. EEXTT [Bre03]. Effective [SBH07, SBH09]. effectiveness [Bre03]. Efficiency [Han02, Bre03]. Efficient [BJK06, BH06c, Dor95, FMS02, HdMKS12, RM04a, Sch98b, Jen03, MMR93, MO92a, Nik05, AAN09, Gun88b, Ooi90, Ruc96, Sud95]. EG [Smi06]. EG-ICE [Smi06]. EGC [MRU12, SHP^{+}05]. Egham [APP12, GP08a]. Egon [AG09]. EGOV [JSWhT11, SJW^{+}12, WTGA05, WSGA06, WSG07, WSF08, WSJT09, WCJS10, TL02, Tra04a, Tra04b]. EGOVIS [AFGvE10, AFGvE11, KLLP12]. EGOVIS/EDEM [KLLP12]. Egypt [BBT08, EKR10, Ima99, SE10]. EHAWC [Dai07, Kar09, Rob11]. EHCI [BPR05, GHP^{+}08, LN01b]. EHCI-DSVIS [BPR05]. Ehrig [KMO^{+}05]. EI2N [MTH08, MHD09, MDH10c, MDH11a]. Eigensystem [SBD^{+}76, GBDM77, Smi74]. eighth [EK81b, NK88]. Eilat [AHW07, DSS^{+}07, ES00b, KS08a]. Ein-Gedi [BtHRBB12]. eInclusion [Ste11e]. Eindhoven [AvLR91a, AvLR91b, ACJ09, BM99a, BLPW03, BLS12, CD09, DN04, DG10, GO09, IdKM06, MBS11, MEAC04, MdRIR08, NQR09, NBM09, ORS89a, ORS89b, Rau04b, RSvdP05, VP96, dBNT87a, dBNT87b, dBBHL10, vES10, vKS12, vdAB03, vdAHW03]. EIS [GHP^{+}08].

EISPACK [Smi74, SBD^{+}76, GBDM77]. Ekaterinburg [DVV07]. EKAW [CP10c, GE08, SS06c, tTVH^{+}12, DC00, FS99a, GPB02, MSS04, PB97, SOS96, SSv94]. EKAW’92 [Wet92]. EKAW’93 [Aus93]. Elba [FC93, MN12a, MN12b]. Elche [Kei09]. Elderly [AGB^{+}11]. Elections [CJR^{+}10]. Electrical [NS81]. Electronic [BGPW05, CJR^{+}10, FPRAW04, FS07a, HAB98, KM04, KLLP12, LSJ06, LPSW04, MZ98, Vei03, Wim03, Wim04, DS01b, DC01, FBSMS05, FRA05a, KW98, LM98a, MSY00, NS99, PSP^{+}02, AY96, AFGvE10, AFGvE11, BW03a, BMP00, BMP01, FMW01, GP89, HAB98, JSWhT11, KYT01, MT09a, SJW^{+}12, TMG10, TMdB11, TMS12, TL02, Tra04a, Tra04b, WTGA05, WSGA06, WSG07, WSF08, WSJT09, WCJS10]. Electrophysiological [CPR^{+}10]. electrophysiology [Sac04]. Elliottville [CM05a]. Elounda [Cou93a]. ELP [Dyc94, DHSH96, LM93]. ELSNET [RG03]. Elspeet [Bod03]. Embedded [BGR08, CH04, CG09, CNmWH^{+}05, DKS^{+}07, DsSL^{+}07, DTV02, GKL^{+}11, GM06b, JQ04, KKP02, KS07a, Kra03, LN09, MS10a, MPPU11, ONPR07, OR08, PV07, Pas02, PFD^{+}10, Pfl02, PT11, PS12c, RS05, Sch04a, SEO^{+}09, Ste07a, SDK^{+}08, Ste09a, Ste11a, Ste11b, WKP03, ZSY^{+}06, ABGP05, DM01a, KP99, KP01b, KNP01, MB98, RV98, AL03a, BDW08, BDSW09, BS05, EYX^{+}05, HPTV05, HK01, KBD87, KSG^{+}07, LKK^{+}07, SVS02, SHX^{+}06, VWH06, VBH07, WCGB05, YGGJ04, YZZ^{+}05, YAL^{+}05]. Embodied [EKF12, KW10, WK08, IPSK04]. Embracing [YAKM^{+}12]. Emergence [VSTN06]. Emergency [LK07, LK09b]. Emergent [SC09a, TWSN07, WAW01]. Emerging [BJSW09, SPSV^{+}04, Paz97, DsSL^{+}07, HJL^{+}09a, HJL^{+}09b, LCDT08, Mue06, Nie09, WZH^{+}07, ZSY^{+}06]. Emeritus [HHG11]. Emile [CA06]. Emilia [SC09a]. Emirates [SSO^{+}11]. EMMCVPR [BKLS11, CBBS09, RFZ03, RVY05, YZCW07, FZJ01, HP99b]. EMMCVPR’97 [PH97]. EMO [CAZ05, EFG^{+}09, ODP^{+}07, TDWG11, Fon03, Zit01]. Emotion [PB08a]. emphasis [Bos96]. Empirical [CW03, FBHW00, KT10, vdADO00, BRS^{+}07, KT10, MN12a]. EMSOA [CW11]. EMSOFT [AL03a, HK01, SVS02]. Enabling [HTMC09]. Enactment [EVV^{+}11]. Encryption [Can12, DR02a, Dun09, GH05, HI10, Joh03, Jou11, Nyb08, Rob06, RM04b, And94, Bih97, Bir07, DRS05, Gol96, Knu99a, Mat02, Pre95, Sch01a, Vau98]. End [CDFP11, PRdRW09]. End-User [CDFP11, PRdRW09]. Enduser [BZ83]. ENEI [BH06d, ED06a]. Energy [LFJI10a, LFJI10b, LJS^{+}10, BKLS11, CBBS09, FZJ01, HP99b, HdMKS12, Leh11, PH97, RFZ03, RVY05, YZCW07]. Energy-Aware [Leh11]. Engineering [AK10, ABB12, APR^{+}06, AMS06, ABBC10, AE05, AD06, AT12, AHJ07a, AHJ07b, AHJ07c, AS07a, AS10, AS12, ADM^{+}12, ABW10, ADP11, Avg12, BMS^{+}08a, BH06b, BFH07, BPS11, BRS^{+}07, BCT02b, BL08, BP11, BCG^{+}07, BCKR10, BCP12, BFGL10, BF11, BW10b, BCSW10, BC08, BHL11, BBPP09, Bra09a, BTT12, BKZ05, BC09b, BHLP07, CSW06, Cer05, CUDL02, CSR08, Cha09, CW09, Che99a, CTS10, CdLG^{+}09, CHCA09, CBS^{+}11, CJ05, CJ06, CGG^{+}07, CHA03b, CP10c, CÓ11, CDSS07, CW03, CYBF06, CY09, CSSW04, COB^{+}08, DF10, DSB04, DF09, DGJ^{+}05, DS11b, DW03b, DZ10b, DMV01, DP06, DRSS01, DL07, EB04, EK09, Ehr85a, EDD04, EOSW07, ELS^{+}10, FS07b, FLVS11, FI08, FP08, FKBA12, FdS12, FKLY08, FTC^{+}10, GHJ06a, GHJ06b, GHJ06c, GGD09, GE08, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c]. Engineering [GLZ^{+}03, GCL^{+}06, GLV09, GTP12, GM89, Gho10, GO11, Gie08, GKL^{+}11, GMO03, GGS11, GH09, GL74a, GL74b, GPB02, GHC^{+}06, GCC^{+}02a, GRP10, GAR04, GRR05, GS06, GB07, HK12, HZK08, Hei12, HA02, Hol07, Hol08, HM09, HS11, HT06b, HKM^{+}07, HKM^{+}11, Ima99, IJ06, JDAW12, KA12, KJ09, KV04, KHJ05a, KHJ05b, KHJ05c, Kie12, KFW04, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KBČ^{+}12, Kou11, KHK^{+}08, KVMH09, KOS07, Küh07, LSV06, LVS08, LLW06, LSgL^{+}06, LHH10, LwCM00, LPH09, Lin01, LCY03, LH06, LMA08, LW12, LPA00, LHJ08a, LHJ08b, LHJ08c, LGR^{+}04, LP08, LGS09, Luo04, Luo06, Luo07, Luo08, Luo09, Luo10, Luo11, Luo12, LDC^{+}09, MBS11, MMO^{+}11a, MMO^{+}11b, MJ07, MNW08, MN12a, MSA99b, MÁ05a, MÁ07b, MVA01, MS02c, MR11, MHS10, MZ06a, Mül10, NM08, NHJ04a]. Engineering [NHJ04b, NHJ04c, NKN^{+}05, NBGA08, NS87, NWHR06, OA07, OYA04, PZ07a, PR08, PHJ03b, PHJ03a, PP12a, PPPM10, Per10b, PS04b, PRH10a, PRH10b, PS03, PMWO02, QQ11, RFBW12, RLL04, RD12a, RT10, RB00, SPH07, SCHS07, Sch01b, Sch03c, SS09b, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, SA12, Smi06, SSZ12, TJ08, Tho11, Tro11, VS09, VRHJ09a, VRHJ09b, VLY09, WMS04, WHG07, WG11, WCK11, WP10, WFB12, WKB04, WH11, Wol07, XL11, YEY04, YAF^{+}02, YTC^{+}07, YWRS11, YCB12, ZS07, ZZZL12, ZSP^{+}04, dLZ12, tTVH^{+}12, vEGW09, vF91, vdBGG10, vdL04, ABS91, Bau75, BMH05, BR92, Bos01, BW05b, BFMS01, BR98, BS01c, BJ01d, CGLR05, CW01, CPvdH03, CMV96, Cue03, CE00, Dei90, DH94, DF96, DC00, DMWW88, DGN01, DMW96, DLW06, EM03, EHH12]. engineering [Enc80, Est98, FRA05a, Fin99, For88, GHM05, GB99, GM03, Gib89, GOW03, GL05, GM99, Got83, GAY86, GN90, GM05, GAR03, Gue05, HCS^{+}05, Hus97, Hus01, Ibr95, ILR95, JO99, JS97, JKB04, JHC02, KHJ05d, Kun84, LB05, LM96, Lon90, Lou92, Lou94, LG05, Luo05, MB82, Mai00, MS03a, MdPA98, MD01, NW01, NL99, OGM05, OP97, PF05, PT98, Pez03, RBC93, RBS93, ST96, SB95, SSW92, Sle92, Smi98, SP93, SSB90, TC95, Tom91, VZ99, WB00, WK02, WBW94, WWC02, X^{+}12a, dPMA98, vdL02, ADP09, AOSV08, APV09, BLS12, BPR05, BSA05, BSH06, BHJY07, CST00, DGR06, ET01b, EWZ11, FP08, GTA10, GOZ05, GHP^{+}08, HTW02, Har07, Har09, Har11a, KOFJ09, LN01b, MRZ09, MWZ10, MVA01, MSS04, OROD07, OTZ01, OPT01]. Engineering [OPP04, PTZ03, SKRZ04, SBH07, SBH09, WBD08]. England [BD98, Fit84, KFLS07, Pat90, Sie86, vH83]. Enhanced [CDS09, DS08, DKW07, KGG^{+}11, NT06, RLKHL12, WKS^{+}10, Kah03]. Enhancement [OYM07]. Enhancing [AH10, BG07a, BG08, DM06, DG06a, Din03, FHH11, FHW12, GA09, MS05, DS03, Fed01, FH01, NA97]. ENQOIR [CLZ^{+}09]. Enschede [Bri97, ERRS10, HLMW06, PHR09, PvS07, SP09, SvS00, vSN01, vdPW10, vdBHOS09]. Entailment [QCDM06]. Enterprise [DW06, DW07b, JNS05a, TQYX11]. Enterprises [ABD^{+}12, BCDN07]. Entertainment [GSH^{+}04, GMI06, cHPkC^{+}07, NRR12, PAD^{+}06, PZE^{+}08, MSW05, AFG^{+}11, HRC06, HMM12, KKKN05, MRN07, ND09, Rau04b, SS09d, YMHH10, ZZP^{+}10]. Entities [ENO12, Sch10a]. Entity [EKT94, FK09, Lou94, PT92, Gog94, Pap95, EKT94, Lou94, PT92]. Entity-Relationship [EKT94, Lou94, PT92, Gog94, Pap95, EKT94, Lou94, PT92]. Entscheidungsproblemen [SS76]. Environment [DMV01, LFJI10a, LFJI10b, LJS^{+}10, WBD08, NY83, Sch93a, WPM05]. Environment-Mediated [WBD08]. Environmental [DT07, Per05]. Environments [BFvdM10, DMV01, GNTF02, Jac07b, Jac11b, Jac11c, KEM07, LMMN10, LMCH12, MTT98, Pol01, Pri03b, SS07b, SS09c, SP07, Ste09d, SGD09, TOL08, vdMBD08, COT98, CDW86, EW91, HNR05, Ish97, KvG98, Lon90, MTO99, Mue01, Nag96, Wei97, SKM07, WPM06, WPM07, Akm87, Bre91]. Environnments [FK09]. Envisioning [Whi00]. EP [HAB98]. EP97 [Ang97]. EP98 [Por98]. ePart [MT09a, TMG10, TMdB11, TMS12]. EPCE [Har07, Har09, Har11a]. EPEW [ABBC10, Bra09a, BKZ05, HT06b, TJ08, Tho11, Wol07, NMP^{+}04]. EPIA [AP11, BCD05, LLMR09, NSM07, BPP91, BJ01c, CC97b, FD93, MPF95, MM89, PA03b]. EPIA’99 [BA99]. Epistemological [Squ09]. EPOS [Squ09]. equational [Hol89, MP96]. equations [AP93, Chi79, Sch92, ZWS81, O’D77]. Equivalence [Thi00b]. ER’97 [EG97]. ER’99 [ABCWM99, Che99a]. Era [CZG^{+}08, CGKM11, EdS08, Paz03]. ERCIM [AFR^{+}04, ABFR07, FRS08, FPFR05, HCF06, LO11, OFR09, OD12b, SP07, Apt00, CS03, O’S03b, SS04]. ERCIM/COLOGNET [HCF06, AFR^{+}04, FPFR05, O’S03b]. ERCIM/Compulog [Apt00]. ERCIM/EWICS [OD12b]. Erfurt [AMU03, Bos01, BJ01d, Cha03a, CE00, FTAZ06, JZ03, JKB04, KOK04, KMTU03, LDTU04, PS03, RMS^{+}05, SKMT04, SS08b, WL04, ZJ04b]. Ergonomics [Har07, Har09, Har11a, vdV84, Dai07, Kar09, Rob11]. Erice [ÉEMS04, MSY08]. Erich [HNR05]. Eringerfeld [BHR90]. Erlangen [HS74]. Erlanger [Ley01]. Error [BS01b, BL07b, BAH09, FILP06, PVW10, Pfl02, BC88, Cal86, CMM93, CGM95, Fos99, FHP03, Gur04, HP89, MMR91, Mor89a, MM97, Pol86, Sak91, CW94b, Don89]. Error-Correcting [BS01b, BL07b, BAH09, FILP06, BC88, Cal86, CMM93, CGM95, Fos99, FHP03, Gur04, HP89, MMR91, Mor89a, MM97, Pol86, Sak91]. Ershov [BBP96, BBZ00, BBZ01, BZ03, CVV12, PVV10, VV07]. Erste [BI73]. ESA [AHW07, AE06, BL05c, DH11, EF12, FS09a, HM08, Nes99, dBM10a, dBM10b, AR04, Bil98, BW97, DZ03, DS96, Len93, Mey01, MR02, Pat00, Spi95, vL94a]. ESAS [BGW06, CHPW05, MTW05, SMCM07]. ESAT [PGV93]. ESAW [ADP09, AOSV08, APV09, DGR06, GOZ05, OROD07, OTZ01, OPT01, OPP04, PTZ03]. Esbjerg [Nür03, OALZ^{+}08, Wii05a]. ESCAPE [CPZ07]. ESEC [NL99, GM89, JS97, NS87, SB95, SP93, vF91]. ESEC/FSE [NL99, JS97]. ESERNET [CW03]. ESF [HAB02a]. Esko [EMO10]. ESO [DsSL^{+}07, ZSY^{+}06]. ESOA [BSH06, BHJY07]. ESOCC [DPZ12]. ESOP [Bar11, Cas09a, De 07b, Dro08, Gor10, Sag05, Sei12, Ses06, Deg03, Gan88, Han98, Jon90, KB92, Le 02, Nie96, RW86, San01, San94, Sch04b, Smo00]. ESOP’99 [Swi99]. ESORICS [AD11, BN09a, BL07a, FYM12, GMS06, GPT10, JL08, dSG05, Ber96, Cup00, DEQ92, Gol94, GKW02, Qui98, SRG04, SG03]. Espoo [AJN12, EHM11, FHJ^{+}02, GU95, HDGK11a, HDGK11b, KSUZ11, LH11, Nyb98, RWD09]. ESPRIT [MS99a, KW96, MT99c, vdL98]. Essays [AG09, ADM11, AAN09, ACÇ^{+}12, ADR08, BS11b, BDR10, BDFM12, BKT12, BCGY09, CGMQ08, CLKK07, CS12b, DHS10, DDM08, Der04, DKS12, EMO10, ELS^{+}10, ELLP12, FJM06, GRS06, GKKW09, HL05, Hei12, JLW07a, JL11, JPR04, KS02a, KS02b, KMPR04, KK11c, KSW03, KMO^{+}05, KR11, Len04, LLM09, LIBP07, MP10a, MM09a, MSS02, Nac12, OKL04, Pal09, PQR11, RSB07, SS06e, BEKS02, HNR05, HS05b, MvOvRdV05, MRS97]. Essen [BF11, CGB11, MCER10, RD12a, WP10].

Essence [MSS02]. Essex [Rai03, RSK03]. ESSIR [ACP01]. ESSLLI [BG12a, IM10, LS12a, NvRSS11]. ESSLLI’97 [Vää99]. ESSoS [BLS12, EWZ11, MRZ09, MWZ10]. ESTAL [VMBMN04]. Este [GPV00]. Estes [CH98]. estimation [CKPB07]. Estonia [BGK12, Bir12, BP93, DG12, DV07a, Dur12, EHKP05, FK12a, GL05, HK98a, JV06, KL12, Lau12, Nie05, O’B12, Pat11, PS12a, PVTG07, Sei12, Uus06, VU05, dLZ12]. Estoril [BSvRW09, BDG^{+}12b, DS09a, FBSMS08, FLL03, HPS09, HMBD09, KHK^{+}08, LGS09, MdlBP05a, MdlBP05b, RM09, VFC08, VASP09]. eSTREAM [RB08]. ESWC [AGS^{+}11b, AGS^{+}11a, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, BHHK08, FKM07, GCFA12, SCP^{+}12]. ESWC2005 [GPE05]. ESWS [BDFS04]. ETAPS [AL11a, AI06, Ama08a, BH06b, Bar11, Bir12, Bod05, Cas09a, Cer05, CW09, De 07b, DG12, Dro08, Dur12, DL07, EM10, FI08, FK12a, GO11, Gor10, GH07b, Gup10, HZ05, Hen08, HP06, Hof11, Kno11, KP09, KO07b, MZ06b, O’B12, Ölv10, Ong10, PS12a, RR08, RT10, Sag05, Sas05, Sei07, Sei12, Ses06, Tho99, dA09, dLZ12, dMS09, Deg03, Due04, GH03, Gor03, GS00b, Han98, Hed03, HM01, Hor02, Hus01, JP04a, KS02c, Kos98, Le 02, Mai00, MY01, NE02, Pez03, San01, Sch04b, Smo00, Ste98, Tiu00, Wal04b, Wat00, WK02, WMS04, Wil01a]. ETAPS’98 [Est98, Niv98]. ETAPS’99 [Cle99, Fin99, Jäh99, Swi99]. Etelsen [MK12]. Eternal [BI11]. ETH [Eng81]. ETheCoM [HP09]. Ethical [PPCS12]. Ethics [KZ10a, KZ10a]. ETICS [BBK12]. Etienne [APV09, TST09]. ETM [SHS10]. ETRICS [Mue06]. ETVC [Nie09]. EU [TAS05]. EUC [DsSL^{+}07, EYX^{+}05, KSG^{+}07, SHX^{+}06, YAL^{+}05, ZSY^{+}06, YGGJ04]. euCognition [EHMM09]. EUD [CDFP11, PRdRW09]. Eugene [MCdS^{+}08]. EUMAS [CKTW12].

EUNICE

[AK10, Leh11, OS09, PvS07, SV12].

EUNICE/IFIP [AK10, SV12]. EurAsia [ST02]. EurAsia-ICT [ST02]. EurAsian [ST02]. EURASIP [AW90]. Eureka [JRR03]. EURO [aCF06, CM05b, GVCA07, KS05b, AC09, ABD^{+}99, BLKW01, BHK01, DGT10a, DGT10b, DLV04, JNR11a, JNR11b, KPS12, KBP07, KBH03, LGG97, LMB08, NWL06, NQR09, SEL09, VS09, CM05b, HAM95, ADB^{+}12a, ADB^{+}12b, ABD^{+}99, BLKW01, BFMR96a, BFMR96b, BFT^{+}08, CAS^{+}09b, DGT10a, DGT10b, DLV04, GVT^{+}11, JNR11a, JNR11b, KPS12, KBP07, KBH03, LMSS07, LGG97, LAF^{+}10, LMB08, MF02, NWL06, PR98b, Sak01, SEL09]. Euro-NF [AC09, NQR09, VS09]. EURO-NGI [aCF06, GVCA07, KS05b]. EURO-PAR [CM05b, ABD^{+}99, BLKW01, DGT10a, DGT10b, DLV04, JNR11a, JNR11b, KPS12, KBP07, KBH03, LGG97, LMB08, NWL06, SEL09, CM05b, HAM95, ADB^{+}12a, ADB^{+}12b, BLKW01, BFMR96a, BFMR96b, BFT^{+}08, CAS^{+}09b, DGT10a, DGT10b, DLV04, GVT^{+}11, JNR11a, JNR11b, KPS12, KBP07, KBH03, LMSS07, LAF^{+}10, LMB08, MF02, NWL06, PR98b, Sak01, SEL09]. Euro-Par’96 [BFMR96a, BFMR96b]. Euro-Par’97 [LGG97]. Euro-Par’98 [PR98b]. Euro-Par’99 [ABD^{+}99]. EUROCAL [Buc85, Cav85, vH83, Dav89]. EUROCAM’82 [Cal82]. EUROCAST [DPA05, DPA07, MDPQA09, MDPQA12a, MDPQA12b, PMD90, MDP04, MDBF01, PMD92, PMD94, PMDA96, PMD97]. EUROCAST’99 [PMDK00]. Eurocode [CC91]. EuroCOLT [FS99b, Vit95, HW01]. EuroCOLT’97 [BD97a]. EUROCRYPT [BCI85, Cra05, Gil10, GQ95, Jou09, Nao07, OF15, Pat11, PJ12, Rue93, Sma08, Vau06, Pic86, Bih03, CC04, CP88, Dam91, Dav91, De 95, Fum97, Gün88a, Hel94, Knu02, Mau96, MZ98, Nyb98, Pfi01, Pre00, QV90, Ste99].

EuroEcTel [PVC^{+}99]. EuroFGI [CA08, CT07]. EuroGP [CTE^{+}06, EOE^{+}07, EAES^{+}10, KTC^{+}05, MSK^{+}12, OVG^{+}08, Pol99a, SFN^{+}11, VGM^{+}09, Fos02, KOL^{+}04, Mil01, Pol00, RSK03]. EuroGP’98 [Ban98]. EuroHaptics [Fer08, IS12a, IS12b, KvETvdH10b, KvETvdH10a]. EuroISI [OALZ^{+}08]. EUROITV [TOL08, CCJ07]. EuroMed [IFGH10, IFL^{+}12]. EuroMPI [CDND11, KGRD10, TBD12]. EuroNGI [CA08, dMS06]. EuroNGI/EuroFGI [CA08]. EUROPE [LCHS96, AvLR91a, AvLR91b, AK07, ASLM09, AGM03, BSW95a, BGP11, BBLT06, BDL08, Bjo87, BHL90, BMJ88, BRW93, BP12b, BAWH00, Chr91, CLT05, CLS07b, CDL12, DKB03, DKN02, EL02a, ES92, FLMG10, FJL97, FHT05, GW96, Gau93, Hal94, HHW01, JZ03, Kat92, Ker97, KV08, KK09b, LS04b, LNSS11, NDB94, ORS89a, ORS89b, OZ01, PH06, PT91a, PT91b, QBA01, RV10, RV11, Slo99, Tou94, Tou96, VW05, WL93, dBNT87a, dBNT87b, Asp98, BS02b, CLS11, CS01b, Gau93, GdlP99, HB97, LPSW04, Lub95, PK00, RS03, Str96, Tou96]. European [ACvLdR03, ACdR^{+}08, AL11a, AV99, ABMM07, ALS^{+}11, AI06, AZ96, ACP01, AT02a, ABK^{+}09, AVP07, AM97, AR04, ABBC10, AD98, APO96, AL04, All96, ARC^{+}07, ABMH08, Ama08a, ACR07, Amb01, Ame91a, AHW07, AL12, ATC^{+}09, AAR05, AD11, AL06, ACJ08, Aus93, AE06, BG10a, BN09a, BYdVZ^{+}12, BBGS10a, BBGS10b, BBGS10c, Ban98, BCD^{+}03, BH06b, BMPM09, Bar11, BHHK08, BD97a, BDG04, BB01a, BHM^{+}04, BdR94, Ber96, Ber00, BB89, BHCL87, Bil98, BD98, Bir12, BL07a, BF00, Bla05, BP01, Bod91, Bod96, Bod05, Bv97, BTF02, BB00a, BBMSD09, BEGP04, BEG04, BRM06, Bra09a, BKZ05, Bra93, BL05c, BDW97, Buc85, BGMST09a, BGMST09b, BW97, BLVV07, BN98a, BN98b, BDFS04, BMP96, BGW06, BC96b, BC96c, Cal82, CL11, CFB^{+}05, CE01, CHD07, Car03].

European

[Cas09a, CW94a, CM95, CHPW05, CIT11, Cav85, CCJ07, CZ05, CW09, CDCJL08, CKM93, Cle99, CFG^{+}11, Coi96, CTE^{+}06, Con00, CS01a, CKTW12, CD01, CDND11, CvH07, CC09b, CM10, CP02, CDS09, CGvR02, CGB11, Cup00, D’H10b, DGM08a, DGM08b, DKP05, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, Dav89, De 07b, DPZ12, Deg03, DG12, DH11, Der92, Des86, DEQ92, DZ03, DKD05, DY10, DS96, Dik04, DS08, Din04, DdCF98, DLM99, DKP00, DLO03, Dro08, Dro09, Due04, DL81, DKW07, DL07, EOE^{+}07, EHP94, ERW99, EHKP05, Ekl94a, Ekl94b, EMT02a, EMT02b, EW91, EF12, Ern07, EAES^{+}10, EM10, Est98, Fau90, FM97, FS99a, FI08, FS09a, Fin99, FS99b, FvdHKL06, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c]. European [F^{+}12d, FLP^{+}12d, FDC12a, FDC12b, FK12a, Fle02, FMN99, Fon00, FYM12, FTZ08a, FTZ08b, FTZ08c, FTZ08d, Fos02, FKBP11, FKM07, FC93, FCLG04, FK95b, FG04b, FSS06a, FSS06b, GCB^{+}05, Gan88, GH03, GB99, GM89, GVB12, GO11, GLM^{+}08, GN88, God05b, Gol94, GKW02, GMS06, GPE05, GTVC06, Gor03, Gor10, GR06a, GS00b, GPT10, Gru98, GO06, GH07b, GMV09b, GKMN10, Gue99, GHMV11a, GHMV11b, GHMV11c, Gup10, GHK^{+}10, HZ05, HM08, Han98, HWP01, HLR^{+}98, HM12a, HK05a, HKM95, HLMW06, Hed03, HL04, HW01, Hen08, HP06, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, HMP99, HSS96, Hof11, HLW08, HM01, HSL^{+}99, Hor02, Hor06, HT06b, HPSZ08, HPZ10, HP99c, HM98, Hus01, HS98d, HKM^{+}11, INR07, Jäh99, JL08, JN10, JS97, JP04a, Jon90]. European [JTB^{+}05, Jul98, KMTW03, KKS11a, KKS11b, KWW04, KS02c, KGL04, KOL^{+}04, KTC^{+}05, KS01, KGRD10, KGG^{+}11, Kno11, Knu01, KvG98, KS04b, KKdM^{+}07a, KKdM^{+}07b, KZ97, KFLS07, Kos98, KFM07, KP09, KKDV02, KKD04, KB92, KO07b, KS91, KMN94, KSGT10, LMR^{+}06, LJR^{+}10, LV07a, LKD08, LW95, LGTB03, LGBT03, Le 02, Len93, LBP06a, LBP06b, LBP06c, LBP06d, LG99, LMV98, Liu11, LB78, LSF07, LP00, LFL05, LMŠT01, LSW08, LKK10, MPP94, MOP^{+}08, Mad92, Mag02, MPP08, Mai00, MN02a, MPS06, MMR07, MM08, MY01, MMP03, MEAC04, MW11, MP10b, MT04b, Mel07, MH11, MNW08, Mey01, Mez11, Mil01, MMK08, MTWD06, MR02, MTW05, Mon96, MSK^{+}12, Mor95, MO05b, MBF08, MHW12, MZF94, MSS05b, MZ06b, NR98, NT06, Nes99, NS87, NE02]. European [NZ03, Nie96, Nie93, NL99, NS98, Niv98, Nob12, O’B12, OVG^{+}08, Ode04, OS09, Olt95, Ong10, Oqu03, OWM04, Oqu07, OALZ^{+}08, PHR09, PM04a, PM04b, PM04c, PM04d, Pat00, PW92a, PM96, PT97, PNPP12, PVP05, Pez03, PH12, PRG09, PRG10, PRG11, PB97, PVC^{+}99, Pol99a, Pol00, Pol01, PvS07, Qui98, RG05, RR08, RCT05, RLKHL12, RW06, RAM05, RRM06, RBTD04, RW86, RS09b, RLT^{+}08, RKM06, RWD09, RT10, RBGG06, RV98, RSK03, Sag05, SRG04, Sam76, San92, San01, San94, SH12a, Sas05, SB95, Sch95, SH08, SDG^{+}07, Sch04b, SJDT08, SPZ90, Seb03, Sei07, Sei12, Ses06, SOS96, SKB10, SFN^{+}11, SHP^{+}05, SF96, Smo00, SC98, SG03, SP93, SC09b, Spi95, SMCM07, SA97, SSv94, Ste98, Swi99, SSZ12, Tho99, Tho06]. European [TJ08, Tho11, Tiu00, TP94, TBD12, TOL08, TdRM^{+}09, TNGK08, Vää99, VTK^{+}12, VP96, VGM^{+}09, Ver08, Ver00a, Ver00b, VDD^{+}11, Vit95, Vit08, Wal04b, War94a, Wat00, WK02, WMS04, WAR94b, Wet92, Wil01a, WKS^{+}10, Wol07, WD00, ZJ04b, ZKZ00, ZHP12, ZQ98, ŻR99, dA09, dBM10a, dBM10b, dSG05, dLZ12, dMS09, dRF01, dRS01, dCHM12, vE91, vHC08, vH83, vF91, vL94a, vW97, vdV84, DS01b, Her91, OAdGBP00, Omi00, PPF^{+}00, ŠPTM00]. European-US [MZF94]. EUROPKI [AL06, CL11, CZ05, LSF07, MP10b, MMK08, PNPP12, KGL04]. EuroPVM [Bod96]. EUROSAM [Ng79, Fit84]. Eurospace [Tou94, Tou96]. Eurospace-Ada-Europe [Tou94]. EuroSPI [ABMM07, Din04, RAM05, RRM06]. EuroSSC [BMPM09, HLMW06, KFLS07, LKK10, RLT^{+}08]. EUSAI [ACvLdR03, MEAC04]. Evaluating [BB95, QCDM06, PDF^{+}09, SG96]. Evaluation [AFP^{+}10, BH07, BP95, CT02, CFH^{+}12, DGT96, FGK^{+}11, FLMK06, FLT07, FKLT08, GKT09, GKT10, GKST11, GKS12, HLR00, HT06b, HHG11, KS09a, KGS08, MCER10, NP09, NP11a, NP12a, PBGK02, PBGK03, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, Sch12, SG07, SBF08, Wol07, BB95, CS87, DN93, Fie02, FLMS05, HK94, HMT99, HBS00, MR97a, ON80, Pet01, PGBK04, PSS98, RMV96, KT10, PBGK02, PCG^{+}07]. Even [GRS06]. Event [Ölv10, SPG10, Dur12]. Event-Based [SPG10]. Events [Bru06b, PJC^{+}12, SKP07, SG07]. Everyday [CA06]. Everywhere [CK07]. Evian [Kah79]. EvoApplications [DCC^{+}10, DBD^{+}10a, DCC^{+}11, DBD^{+}11, DAC^{+}12]. EvoBIO [GVB12, MMR07, MM08, PRG09, PRG10, PRG11, RCB^{+}04, RBC^{+}05, RBC^{+}06, Rai03]. EvoCOMNET [DBD^{+}10a, DBD^{+}11, DAC^{+}12, GBC^{+}08, GBC^{+}09, RCB^{+}04, RBC^{+}05, RBC^{+}06, Gia07]. EvoCOMPLEX [DCC^{+}10, DCC^{+}11, DAC^{+}12]. EVOCOP [RG05, CvH07, CC09b, CM10, GR06a, HM12a, MH11, vHC08, Boe01, Cag02, Rai03]. EvoENVIRONMENT [DBD^{+}10a, GBC^{+}09]. EvoFIN [DBD^{+}10a, DBD^{+}11, DAC^{+}12, Gia07, GBC^{+}08, GBC^{+}09]. EvoFlight [Boe01, Cag00]. EvoGAMES [DCC^{+}10, DCC^{+}11, DAC^{+}12, GBC^{+}09]. EvoHOT [DBD^{+}11, DAC^{+}12, GBC^{+}08, GBC^{+}09, RCB^{+}04, RBC^{+}05, RBC^{+}06]. EvoIASP [DCC^{+}10, DCC^{+}11, DAC^{+}12, Gia07, GBC^{+}08, GBC^{+}09, PVC^{+}99, RCB^{+}04, RBC^{+}05, RBC^{+}06, Boe01, Cag00, Cag02, Rai03]. EvoINTELLIGENCE [DCC^{+}10, DCC^{+}11]. EvoINTERACTION [Gia07, GBC^{+}09, RBC^{+}06]. EvoLearn [Boe01]. Evolution [BEJvR11, BOH^{+}12, Che99a, CML^{+}10, DBC^{+}10, HLC^{+}12, HKV10, KM11, KD11, LKZ^{+}08, LCF^{+}04, Mit05, Mit06, MS08a, MTC^{+}08, OLL10, PST11, VSTN06, WLC^{+}06, All96, BdBC01, BEM91, Col02, Del00, Fon00, HLR^{+}98, McK99, PGV93, TLC06, YKF97, vdL98, BE95, Sip97]. Evolutionary [BEM91, BDMS08, CPFD^{+}03a, CPFD^{+}03b, DSM94, DPB^{+}04a, DPB^{+}04b, DCC^{+}10, DBD^{+}10a, DCC^{+}11, DBD^{+}11, DAC^{+}12, EV96, Gia07, GBC^{+}08, GBC^{+}09, PDSD10, PSDS11a, PSDS11b, PDSN12, RCB^{+}04, RBC^{+}06, RKS^{+}12c, Boe01, Cag00, Cag02, FU96, Rai03, RBC^{+}05, ST96, Yao95, Ang97, CAZ05, CS97, CvH07, CC09b, CM10, EFG^{+}09, Fog94, Fog95, Fog96, Fon03, GVB12, Gom01, GR04, GR06a, HM12a, HM98, MRC12, MMR07, MM08, MH11, ODP^{+}07, Pal01, PRG09, PRG10, PRG11, PVC^{+}99, Por98, RG05, TDWG11, Zit01, vHC08]. evolvable [ST96, HML97, HSH08, KLZ07b, Liu01, Mil00, MMC05, SMPU98, TTM10, THT03]. Evolving [MdSC09]. EvoMUSART [DBD^{+}10a, DBD^{+}11, Gia07, GBC^{+}08, GBC^{+}09, MRC12, RCB^{+}04, RBC^{+}05, RBC^{+}06, Rai03]. EvoNUM [DCC^{+}10, DCC^{+}11, DAC^{+}12, GBC^{+}08, GBC^{+}09]. EvoPAR [DAC^{+}12]. EvoPLAN [Cag02]. Evora [APO96, BA99, AS11a]. EvoRISK [DAC^{+}12]. EvoROB [Rai03, Cag00]. EvoRobot’98 [HM98]. EvoSCONDI [Cag00]. EvoSTIM [DBD^{+}11, DAC^{+}12, Boe01, Cag00, Cag02, Rai03].

EvoSTIM/EvoPLAN [Cag02]. EvoSTOC [DCC^{+}10, DCC^{+}11, DAC^{+}12, Gia07, GBC^{+}08, GBC^{+}09, RCB^{+}04, RBC^{+}06, RBC^{+}05]. EvoTel [Cag00]. EvoTRANSLOG [DBD^{+}10a, DBD^{+}11, GBC^{+}09, Gia07, GBC^{+}08]. EvoWorkshops [Gia07, GBC^{+}08, GBC^{+}09, RBC^{+}05, RBC^{+}06, Boe01, Cag00, Cag02, Rai03, RCB^{+}04]. Evry [BCBF08]. EWCBR [BP01, HKM95, SF96, SC98, WAR94b]. EWCBR-94 [HKM95]. EWCBR-96 [SF96]. EWCBR-98 [SC98]. EWHCI [BUG93, BGU94, BGU95]. EWICS [OD12b]. EWLR [BD98, WD00]. EWLR-6 [BD98]. EWLR-8 [WD00]. EWMF [ABG^{+}06, BHM^{+}04]. EWNLG [AZ96]. EWRL [GLM^{+}08, SH12a]. EWSA [GO06, MO05b, OWM04]. EWSL [Kod91]. EWSL-91 [Kod91]. EWSN [LV07a, MW11, PH12, RS09b, RKM06, SKB10, Ver08, KWW04]. EWSP [Her91]. EWSPT [Amb01, Con00, Der92, Mon96, Oqu03, Sch95, War94a]. EWSPT’98 [Gru98]. Exact [BL06b, CF09a, DFD04, GN08, MR12a, RS10, TW12]. Example [RRMP08]. Excellence [CA08, aCF06, GVCA07, KS05b, TAS05]. Exception [DKRT06, RDKT01]. Executable [FO95]. Execution [Bee00, BG91b, Che92]. Exercises [Eve75]. Exeter [LJ00, YEY04]. Exhibition [Bau99, GH94a, GH94b, HS95b, HS97, LCHS96, SBH98]. Existence [PCH^{+}06]. expanding [FNY93]. Experience [BLW09, BW11, CCJ07, CDLZ06, CMS08a, DW07a, DBPZ05, KDWW10, Mar11a, Mar11b, RSW12b, CS03, KK95, Ber02a]. Experiences [CW03, DKW07, BL94, FNPW05, II00, Neh88, Vaz99]. Experiment [PF07b, Pet80]. Experimental [CPZ07, Rei99, AS06, Dem07, Fes10, FMS02, Jen03, Kla12, McG08, Nik05, PR11, RM04a, RBS93, Vah09]. experimentation [BS01c, GM99]. Experiments [JMP12, LOR10, MW08, MS02c, PDK^{+}10, PCG^{+}10, SW08, BBB98]. Expert [AE05, BKW08, BKW09a, BKW06, BHQ10a, BHQ10b, CHA03b, GTT04, HLSZ11a, HLSZ11b, HKW^{+}12b, HA02, IKR00, Ima99, LLP^{+}07, LSTZ12a, LSTZ12b, LPA00, MRŠ03, MVA01, OYA04, QSBC98, WRP07, ADW05, BR92, BCST99, HT97, HCT02, Kar94, MLW93, May01, MdPA98, RT95, WT96, dPMA98, GN90]. Expert-Systems [HKW^{+}12b]. Explicit [Dad01]. Exploitation [Urs02, Boc95]. Exploration [DKN^{+}11, ABD04]. Explorations [IM10]. Exploratory [HAB02a]. exploring [vLWH^{+}00, YAKMJ11]. explosion [God96]. Expressions [Dad01, EV09]. Extended [AGS^{+}11b, AGS^{+}11a, AAH^{+}10b, AAH^{+}10a, SCP^{+}12, Gog94, YBM87]. extendible [Pau93]. Extending [ABG96, BTT90, BCP^{+}04, ISS^{+}06, JBJ94, JJP^{+}02, PDG92, SSRA98, SCM88, Zan00]. extension [BK90, GBDM77]. extensions [DPP95, DPP97, Dyc94, DHSH96, EHSH92, LM93, Rao95, SH91]. Extracting [SKP07, Sta99]. Extraction [BJ01c, PST11, Paz99, Paz03, Paz97, Urs02]. Extreme [CDSS07, MdSC09, ZEL04, AMS06, BMH05, EB04, MS03a, MW03, WW02, ZEL04]. Ezhou [YZCW07].

FA [AVP07, HK05a, PHR09, RW06, SH08]. FAABS [HRTR05, HRT^{+}03, Ras01]. FAC [SR07, SEG09, SF11]. Faces [ZZTG07, ZGT05]. Fachtagung [BI73, BHH74, Lei74, SF74, Sch74, KF74]. Facial [EV09]. Facility [CHR10, HRdV11, SL12]. Facing [KKW11, KKW12]. FACS [AÖ12, BL12]. factorization [Tha84]. factors [BZ83, LSW81]. fail [Pfi96]. fail-stop [Pfi96]. Failures [Bee00]. FAIR [KJ91]. Fairfax [For88]. fairness [Rao95].

FAlgorithmic [CBNR02]. Fall [Nie09]. Falls [BFvdM10, Die03, EK09, LV02, Ngo09, STY^{+}07, THX^{+}07]. families [vdL98, vdL01]. Family [vdL04, HKB93, vdL02]. FAPR [GO96]. FAPR’97 [Gab97]. Faro [Fon03, HDBM11, MBT^{+}03, RM02]. FASE [BH06b, Cer05, CW09, DL07, FI08, GO11, RT10, dLZ12, BD97b, GJ93, Hus01, Mai00, MNS95, Pez03, WK02, WMS04]. FASE’98 [Est98]. FASE’99 [Fin99]. FASec [ARS03]. FAST [BDE12, DGM09, DG10, DEG11, DMRS06, DMRS07, Pfl02, And94, Bih97, Bir07, Can12, DR02a, Dun09, GH05, Gol96, HI10, Joh03, Jou11, Knu99a, Mat02, Nyb08, Pre95, Rob06, RM04b, Sch01a, Vau98]. faster [CKPB07, Pan84]. FATES [GN05, GW06b, HNRW06, NBM09, PVTG07, SHUH08, PU04]. Fault [BJRT06, BJRT09, DO02, Jos00, MCER10, Sch12, BL94, JP96, Jos88, LY04, LdRV94, ON80, RR98a, Vyt91, BKNS06, SS90]. Fault-Tolerant [BJRT06, DO02, Jos00, JP96, Jos88, LY04, LdRV94, ON80, RR98a, Vyt91, SS90]. FAW [ALZ11, DHX09, LCY10, PF07a, PWY08, SLSW12]. FAW-AAIM [ALZ11, SLSW12]. Fayetteville [DS09b]. FC [BDC12, CSN14, DDS12, Dan12, DR06, DD07, DG09b, Ker12, PY05a, Sad13, Sio10, SCD^{+}10, Tsu08, Wri03, Bla03b, Fra01, Fra99, Hir97, Hir98, Jue04, Syv02]. FCRC [LM96]. FCT [CVÉ07, Gec81, Kar77, Kar83, KCG09, LR05a, OST11, Bud91, CC97a, Esi93, Fre01, LN03b, Rei95]. FCT-Conference [Gec81, Kar77, Kar83]. FCT’87 [BBL87]. FCT’89 [CDG89]. FCT’99 [CP99]. FDTC [BKNS06]. Fe [ASAM07, COT98, FK87]. Feature [SGG06]. Features [EBAH08, Mül07a, RME04, Rei93]. February [Abe06, ABLM05, AW90, AH03, ABD03, ADL09, BLS12, BFBS11, BG11b, Bjo80, BCP94, BvLR04, BBPP09, BVNW87, BUHS08, BS06c, CR05a, CH04, CL90, Cho02a, Cho02b, Con00, CW88, DR02a, DDS12, Dan12, DGT96, DU09, DFM04, DG02, Del00, DR06, DPA05, DPA07, DD05, DD07, DG09b, DKN12, DH06, Dun09, Dun12, DS06, DT06b, ES02, Ekl04, EFW93, EMW94, EW03, EWZ11, FV03, FR01b, FJ92, FILP06, FP05, Fra01, Fra99, FGK^{+}10, GG05b, Gel01, Gel02, Gel03, Gel04, Gel05, Gel06, Gel07, Gel08, Gel11b, Gel11a, GH05, Gol96, GM07b, HKI08, HRB12, Hir97, Hir98, HI10, IZ98, JM10, Joh03, Jou11, Jue04, KS03a, Kah03, KG04, KK97, KK11b, KM08b, KP04, Ker12, Kia11, Kil05, Kim01a, Kim05a, KPS01, KW10, KB92]. February [KS07c, LM93, LLMT04, LW09, Lin14, LP10b, LK07, MWD06, Man75, MW11, Mar03, MRZ09, MWZ10, MO08, MS03c, Men05, MJ07, Mic10, MS06b, MC89, MDP04, MDBF01, MDPQA09, MDPQA12a, MDPQA12b, MMK98, MSKY10, NP02, iNR08, Nao04, NO11, Nie94, Nyb08, Oka04, PS02a, PS97a, PS98a, PY05a, PH12, PMD90, PMD94, PMD97, Pla96, Poi06, Pre02, Pri03a, Pri03b, PR96, RF10, RiN12, RR12, Rao92, RT00, RM97b, RBTD04, RS09b, RKM06, RM04b, RME04, SGG06, ST00, SSW92, STT04, SKB10, SHP^{+}05, SK08, SKB98, Syv02, TBC07, TW07, Vad07, Ver08, VS05, War94a, Wil99, WHR10, ZVU99, vdL98, vM94]. February/March [Nie94]. Federal [BE87, BVNW87, DMWW88, Nic86, Ott87, RW86, TS87]. Federation [JLST06]. Feedback [DJNR06, MMBND07, MSS05b, Ron84]. Feliu [PB97]. Fellows [BDFM12]. Ferran [MRU12]. Ferrand [CP12, KS07c, RMTW11]. Fertilization [CCM^{+}04]. Festschrift [SW10, TJHS94]. FET [PQ05, Pri03b]. Fethiye [RBGG06]. Feuerstein [Bet83]. FG [dGEK11, dGN12]. FGIT [hKhLKŚ10, hKAS^{+}11, hLhKcFŚ09]. FHIES [LW12]. FHPCN [M^{+}06]. FIDJI [GRR05, GAR03, GAR04]. Field [GH93, HS94, HG96, BPV04, BW01, CCdS03, GZR02, GH93, HS94, HG96, HK98a, HG00, LCG97, LIH99, ML95]. Field-Programmable [GH93, HS94, HG96, CCdS03, HG00, BW01, GZR02, GH93, HS94, HG96, HK98a, LCG97, ML95]. Fields [CS07, HH10, MPS04, ÖRH12, vzGIK08, BDR10]. Fiesole [GS10]. Fifth [AKL^{+}99, Can08, KLN07, Luk97, Müh75, PMDA96, PP04, RM09, AB78, Mah85, Mic90, Sko85, SPSV^{+}04]. Fight [KT11]. FIHM [SS07a]. file [Kis95]. Filtering [NMS93]. FIMH [ADS09, FRSH05, MA11, Mag03]. Fin [GT93]. Final [Ceb08, Eug09, Sex09, Smi03, SC07, TSE^{+}09, AS02b, BFS92, BP95, BJ95, BBC04, EDD04, HR91, JL95, MØ05a, MS99a, MO92a, RMV96]. Finalists [RB08]. Financial [LwCM00, Sad13, Wri03, Bla03b, BDC12, CSN14, DDS12, Dan12, DR06, DD07, DG09b, Fra01, Fra99, Hir97, Hir98, Jue04, Ker12, PY05a, Sio10, SCD^{+}10, Syv02, Tsu08]. fine [Boc95]. fine-grain [Boc95]. Finite [CS07, HH10, ÖRH12, Sch03b, Tau89, vzGIK08, Pin89, Sch97b, Sif90, MPS04, YJKK06, YJKSW10]. Finite-State [YJKK06, YJKSW10]. Finland [AMS06, AM97, ACJ08, AJN12, BBKW03, BO00b, BKS05, BM98, BPS02, CKPR11, CMS08b, DKLS04, DT06a, EMT02a, EMT02b, EHM11, FHJ^{+}02, FKS10, FK12b, GU95, HKL07, HHLS08, HKT12, HDGK11a, HDGK11b, ILR95, IS12a, IS12b, JN10, KPK05, KS12a, KTH09, KMC^{+}01, KSUZ11, KOOL03, KTB09, KC02, LH11, LS88, NR04, NU92, Nyb98, OKHH^{+}08, OKS02, OYY03, Oqu03, PRT06, PS02b, PB08b, RKM^{+}12, RRM06, RYG11, RWD09, SGPP06, SS77, SMK02, YJKK06, vGH96]. FinTAL [SGPP06]. FIRA [HSP^{+}12, KGV^{+}09, PT09]. First [ACvLdR03, AB10, ARS03, AKK04, AST05, AV07, ABB10b, AFGvE10, ACH00, AM98, AJ01a, AKSO06, BB07, BFF^{+}10, BCL03, BMP00, Bel03, BP11, BHM^{+}04, BM99b, BCB97, BJN03, BI73, BFMS08, BBBB08, BGKS04, BDGK00, BH00, BKM06, Buc90, BY06, BDFS04, BCDN07, BLS02, BB00b, CPR^{+}10, CMSM^{+}10, CS05a, CS07, CNP^{+}08, CHPW05, CSW06, CT07, Cha02a, CL06a, CPZ07, CY10, CLZ06, CJ05, CH99, CM06, CH96, CW01, Coh94, Col07, CNmWH^{+}05, CEKR02, Cov99, CRS00, CHR10, CE00, DKY06, DDE04, DSLT08, DMO98, DPZ12, DG02, DG12, DBPZ05, DY06b, DLdS04, DHO08, DFT09, DW94, DFD04, DXZ07, DTF98, DKR02, Duf07, DS06, EKO08, EM05, EBT07, Est98, ER12, FV01, Fau90, FLY06, FdCLJ^{+}11, FdCLN^{+}11, FNY93, Fis08, FLF^{+}08, FCLL02, FKW08, Gab97, GPAS^{+}06]. First [GM01, GS01, GLV09, Gel99, GB77, Gie10, GLN01, GHH06b, Gup99, GM07b, Ham08, HTW02, HAM95, HK05a, HNRW06, HLMW06, HLKX12, HS98a, HK01, Heu98, Hir97, HR08, HW07, HH02, Hor06, HM12c, HKS09, HM98, HMSU04, HdMKS12, IdKM06, IMW04, IZ98, Ish99b, IGN05, IFV07, Ish97, JZR09, JMBS07, JM05, JSO04, JM95, JWH06, JS11b, JMT11, KK97, KWW04, KMC^{+}01, KGL04, KLP04, KP10a, KLP10, Kim01b, KR00b, KP99, KMS12, KT11, KR04b, KMMP12, Kur09, LJS08, LLW06, LN04a, Le 94, LBP00, LV02, hLhKcFŚ09, LOST04, Leo00, LLP^{+}07, LMP04, LYCV08, LSK08, LSIT11, LW12, Llo00, LK07, LOH05, Lop06, Lor01a, Lor01b, LZ00, Luo04, LCDT08, MRC12, MT09a, MPP08, MP06, Man75, MSS11, MPdSM07, Mar11a, Mar11b, MS06a, MMV04, MRZ09, MN02b, MB06]. First [MSS03a, MJ07, MW08, MB09, MMR93, MMS93, MÁ05a, MÁ05b, MSG06, MOT12, MT99b, MS08b, MK03b, MB97, NL01, NP09, Nao04, NJ10, NZ06, NT06, Ngu06, NGHJ07, NKC09, NR04, NK99, Niv98, NT03, NS99, ND04, Oho03, OTZ01, OWM04, Oqu07, OWPY03, OALZ^{+}08, PAD^{+}06, PS97a, PDSD10, POP09, PH11b, PCPS03, PP99, PT97, PK03, PVC^{+}99, PS98b, PS07b, PISW05, PF07a, PLMV12, Pri03a, PS09c, QCDM06, Rah04, Raj09, RLL04, Ras01, RWY97, Ret97, RMS^{+}05, RBTD04, RF99a, RSL01, RH01, RCLE05, RT09, RPZ97, RFBD09, RO00b, SNY06, SGSV10, SMBQ07, ST00, SKMT04, SJDT08, ST12b, SMP07, ST02, SdDK91, SCG98, Smi05, SPSV^{+}04, SS08b, SB12b, SKCK07, SG07, SKB98, SLM07, SHŚ^{+}07, TOOO07, TST10a, TST10b, TBC07, TN04, Tou94]. First [TL02, TGM97, THR03, TtHVS06, TWSN07, UBW06, VS09, VGP08, Vau08, VA95, VC03, Vor92b, VASP09, WYL01, WPSY06, WYSS10, WC98, WW02, Wes01, WMdLA10, Wil99, WNL05, WPZA11, WL93, WCGB05, XHXL08, XLHX08, Yak00, YKF97, YSR^{+}06, ZRP^{+}09, ZL96, ZJ04a, Zha07, ZLY^{+}07a, ZYH03, ZW09, ZZLL10, Zim92, Zit01, dBBGdR03, dMS06, dMGFS11, dRWK^{+}10, ter97, vES10, vdHI99, AJS05, AH94, And96, Aya95, BL05a, Ban98, BBKP06, BB05a, BGD^{+}05, BS94, Bri95, BHK01, BR96b, Che03a, CSV05, Coh92, CLT05, CFP94, DV05, DDFM05, DL95b, EHP94, FHRR05, Fri93, GO94, GM04a, Gue05, HOQ97, HP95, Hee94, HPR05, HML97, HVT12, JKB04, Jou94, Kar91, KKM04, KT96, KZ97, KS05a, KS05b, Len93, LWD05, LR94, LA05, Mac06, MRK05, MSW05, MXZ05, MR95, NY93]. first [OFK05, PBB05, PS96b, Pie05, PB92, RMR05, SSW92, SC96, SLP05, STLK05, TTP05, VWY97, WCO05a, WCO05b, WCO05c, Wat95, WAR94b, WPM05, YRMD05, ZHF04, ZLJ05, Zho01, BMS01, Har79]. First-order [Har79]. FIS [BGPTF10, DFT09, ZSP10]. Fischbachau [BS92, CS00b]. FITraMEn [VS09]. Fitting [Ahn04]. FL [AN09, BGP08, BTMI11, BHLP07, DL04b, Duf11, FK09, Har11a, HW06b, Jac11a, Jac11b, Jac11c, Jac11d, JS02, Kur11, MNZ09, Mar11a, Mar11b, OPS11, OZ11, PM09, RVY05, Rau11, Rob11, SS11c, SF11, Shu11a, Shu11b, SS11d, Ste11e, Ste11f, Ste11d, Ste11c, Baa03, CPvdH03, DG02, Des02, FSM03, Jue04, LN03a, SW03b]. Flagstaff [BG93a]. FLAI [KS93]. FlashDB [XYZU11, YYH^{+}12]. Flexible [ACL98, AMCL02, AYB^{+}09, BH06c, CHAL04, CDY^{+}11, LPOA^{+}06]. FLOPS [BKV10, GH08, HW06a, ST12a, HRA02, KS04a, KU01]. FLOPS’99 [MS99b]. Florence [ACR12, BCST99, Bla80, CVLL12, Del97a, Del97b, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, FL11, FMC12a, FMC12b, FMC12c, GS95, MD04, MT99b, Nes95, Tah01]. Florianopolis [KKM04]. Florida [FR98, JL00, OPS11, PB07, Rol98, SL99a, LSW00]. Flow [MO06]. fluid [Ste12]. FM [BKZ05, BNZ06, BB11, BW09, CP12, DH08, LS10a, BS11f, CD09, CMS08b, FHT05, GM12, HSTU99, MNS06]. FM-Trends [HSTU99]. FM8501 [Hun94]. FM’99 [WWD99a, WWD99b]. FMCAD [HM04, AO02, GW98, HJ00, SC96]. FMCO [AdBB12, dBBGdR05, dBBGdR06, dBBGdR07, dBBGdR08, dBBM09, dBBHL10, dBBGdR03, dBBG04]. FME [AGM03, DB04, EL02a, FJL97, GW96, NDB94, OZ01, WL93]. FMICS [ACJ09, BHLvdP07, CF09b, KR10b, LM08a, SS11b, SP12]. FMN [MZCC09, ZCCL10]. FMOODS [BdB08, BJ07, BD11, GR12, GW06a, HZ10, LLPH09, SZ05, NNS03]. FMSB [Fis08]. Focused [BI04, BKS05, GKT09, GKST11, OKS02, BO00b, BKS01, BVO10, COBV11, DJJ12, FKLT08, GKT10, GKS12, JS08, MV06, MA07c]. Focusing [Rei99]. FODO [LS89, Lom93]. FOGA [STWS07, WVDS05]. FoIKS [DH06, HKI08, LP10b, LS12b, ES02, STT04, ST00]. Folding [Sch03b]. FoMLaDO [BdBC01]. FoMLaDO/DEMM [BdBC01]. FoMLaDO/DEMM-2000 [BdBC01]. Font [CJ95]. Fontainebleau [GPR^{+}12]. Food [PS08a]. FOPARA [PvES12, vES10]. foreign [ZZZL12]. Forensics [GFV09, SFS11, SKPG12, SF08c]. Forestalling [Bee00]. Forever [KMPR04]. Foreword [CA06, Koh06b]. Forging [GCSR07]. form [CAdB01]. Formal [ABDW10, AF99a, AT12, AGM03, BPS11, BI03, BH07, BDZ08, BPZ09, BI11, BCP12, BHHJ11, BI73, BDS04, BDGK00, BEKS02, BKZ05, BNZ06, BB11, BC09b, BW09, BHLP07, BS11f, CP12, CD09, CB06, CMS08b, DSB04, DG07a, DSS11, DB04, DGLT12, DW03b, DZ10b, DKOZ05b, DH08, Ekl04, EL02a, FHT05, FBHW00, GO96, GSW05, GTP12, GM03, GM12, GO09, GP12, GSS00, GKKW09, HKP11, JK06, KMR12, LY04, LS10a, LB05, LSR03, LW09, LH06, LMA08, MS06a, MS10b, MS10c, MS12b, MM10, MNS06, MOT12, NMP^{+}04, OW09, OKL04, PM08, PU04, PT91a, PT91b, QQ11, BC04a, BB05b, Ber02b, BBKW03, BO80, Bjo87, BHL90, BH95, BHT97, BFH98, Bra75, BHK01, BPS02, EHH12, Eng74, FJL97, GW96, Haj96, HSTU99, KBD87, MSS12, NDB94, OZ01]. formal [OK90, PS97b, RSvdP05, TKHS05, WWD99a, WWD99b, WL93, AO02, ARS03, ABL96, ADM11, AM03, AdBB12, ABDW10, ACJ09, AT12, AÖ12, AB06, BL12, BdB08, BDE12, BM11a, BDG12a, BI03, BC04a, BB05b, BH07, BDZ08, BPZ09, BI11, BCP12, BBP93, BJ07, BKZ05, BC09b, BHLvdP07, BMS96, BD11, BHLP07, CJ08a, CH10, CB06, CF09b, CKOS09, DO02, DSB04, DSS11, DGM09, DG10, DEG11, DV07a, DMRS06, DMRS07, DIP12, DW03b, DZ10b, EK09, Ehr85a, FT11, FR09, Fis08, GG05b, GM03, GN12b, GR12, GW98, GW06a, GN05, GW06b, HZ10, HNRW06, HBH08, HRT^{+}03, HRTR05, HT06b, HM04, HJ00, Hus97, JLW07a, JP96, Jos88, Jos00, JN12, Kat99, KHW03, KR10b, KMO^{+}05, KS97, KS07c, KS10c, LY04, LdRV94, LN04a, LB05, LLPH09, LM08a, LL95]. Formal [LH06, LMA08, MO08, MS99a, MS06b, MPS93, NNS03, NPPDG06, OW09, OV09, PV02, PU04, PY05b, Pin89, QQ11, Ras01, RT07, RR98a, SS11b, SM11b, Som09, SC96, SZ05, SP12, SW75, SHYEF08, VJ11, Vyt91, Wan05a, Wol07, dBBGdR03, dBBG04, dBBGdR05, dBBGdR06, dBBGdR07, dBBGdR08, dBBM09, dBBHL10, dFEN04, dGEK11, dGN12]. Formale [BI73]. Formalism [MNW08, Ste12]. formalisms [PWZ98, dBdRR90, HP98]. Formalization [DR81]. Format [Koh06b]. FORMATS [AB06, CJ08a, CH10, FT11, JN12, LY04, OV09, PY05b, RT07, LN04a]. FormEMC [NMP^{+}04]. formerly [HMPZ09]. forms [Nij80, Woo80]. Fort [DL04b, FR99b, LN03a, MNZ09]. Fortaleza [DLdS04, dSDL04]. FORTE [BD11, DV07a, GR12, HZ10, LLPH09, NPPDG06, SHYEF08, Wan05a, KHW03, NMP^{+}04, PV02, dFEN04]. FORTEST [HBH08]. Forum [Ada96, AFP^{+}10, BHM^{+}04, FGK^{+}11, GNR07, OO12, Pet01, PBGK02, PBGK03, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, PRR05, RR01, RR03, RBG09, RV98, PGBK04]. FOSAD [AGM05, AG07, ABG09b, AG11, FG04a]. FOSSACS [AI06, Ama08a, Bir12, Hof11, Ong10, Sas05, Sei07, Tho99, dA09, Gor03, HM01, NE02, Tiu00, Wal04b]. FoSSaCS’98 [Niv98]. Foundation [BF11, EDD04, GH09, PR08, RD12a, SPH07, WP10]. Foundational [AM03, PvES12, vES10]. Foundations [AS02a, AVP07, AS03, AN07, AN09, AKG06, AP07, AKR03, AAR05, BdBC01, Ber02a, BLvdH10, BCGY09, BSG03, Cai05, DSS11, DFP^{+}12, DO89a, DO89b, DMWW88, DR02b, DKS12, EK09, EEK^{+}87, EKLM87, FKK04, FKBP11, GSW05, GN12b, HRK^{+}07, HMV01, HK05a, HK10b, JS97, KP00, KHG01, KU06, KN09b, KK07, KWT03, KSGT10, KR11, LBSBW98, LM04, MSA99a, MS11b, NR00, NL99, OT08, OW09, PHR09, PR03a, PAC08, RS06, RW06, RV03, RSW12a, SST00, Sch98b, SH08, SM11b, Ste01, VTK^{+}12, WZ09, AN97, Bec75, Bec79, Bli75, BS93, Bro92, Bro94, CK84, CJK88, Dem80, Ehr85b, FJL97, FM99, Gru77, GC81, GRW86, HK92, HKNW07, Hus97, JS05, KMS89, KPW99, LSW00, LKZCS05, Mai88, Mai90, Maz76, Mel86, MT89, MT99c, Mon01, NT92, NM94]. foundations [PS96a, PD96, PRR94, PR97, Rov90, SPK01, Tar91, WH95, Wil91, Win78, ZBG98, dBdRR93, van97b, AI06, AGM05, AG07, ABG09b, AG11, Ama08a, AMRŚ08, AR98, Bir12, BN91b, BAL88, BFJV97, CJ11, CV96, CFLR12, CS10, DLFP02, DV09, DH06, ES02, ERMS06, FSdlPT11a, FG01, FG04a, GAL12, Gor03, HRZK02, HMRT05, HKI08, Hof11, Hol89, HM01, JS84, Kar83, KRSR11, LN94, LP10b, LS89, LW12, Lom93, LS12b, Mad89, Mah85, MCA^{+}07, NE02, NCW97, Niv98, Nor86, Nor87, NK88, Nor90, Ong10, RRR99, RS97a, RM96, RS97b, RS99b, RO00a, RRBE09, Sas05, ST00, SR07, SEG09, SF11, Sei07, STT04, Shy92, Shy93, STWS07, Thi94, Thi95, Tho99, Tiu00, Wal04b, WVDS05, ZRTS03, dA09, dBdRR91, BJM^{+}05, BE05]. Founded [Ull08]. four [Kow01]. Fourth [Ant97, BC96b, BC96c, DGSM11b, DS96, DMRS07, EEM96, Fek08, FMRW09, GPYD09, GS95, HL10, KM00, MSZ08, MBC^{+}97, PR00, RABR03, SEO^{+}09, STA^{+}11, Sie75, SPSV^{+}04, SGD09, WLS89, Wol07, YNB08, AZ96, A^{+}08, Ban92, JS84, LMV95, PV04, RKS89, Rob80, SS77, vP93]. FoVeOOS [BM11a, BDG12a]. FP [DMB^{+}11, HRK^{+}07, HP09, SPC^{+}08, TDK^{+}10]. FP-UML [DMB^{+}11, HRK^{+}07, HP09, SPC^{+}08, TDK^{+}10]. FPGAs [HK98a]. FPL [BW01, CCdS03, HG96, BPV04, GZR02, HS94, HG00, LIH99, ML95]. FPL’97 [LCG97]. FPL’98 [HK98a]. FPLE [HP95]. FPS [GAL12]. Fq7 [MPS04].

FQAS [AYB^{+}09, CDY^{+}11, LPOA^{+}06, ACL98, AMCL02, CHAL04]. frames [SK97, SZ00]. Framework [CDE^{+}07, DHM99, Kno98, PBCF08, ZZ04b, Lub95, Naz94, Pfi96, Sin94, Lam00, LSR03]. Frameworks [Pas02]. Franca [GI03]. France [APFV09, ARS09, AP93, ABB12, AK90, AV99, AGK^{+}10, ARSB09, AAA^{+}11, AW12, AF97, ABCWM99, ABFV12, AG98, ADB^{+}12a, ADB^{+}12b, AD06, All96, Alo05, AF02, AC09, AL03b, ABD^{+}99, AFGvE11, ADL05, ALR11, ABG96, APV09, AB06, AJ01a, AJ01b, APM08, Aus93, ACD^{+}10, Aya95, ADS09, ADGM12a, ADGM12b, ADGM12c, AB97, Baa07, BN09a, BCBF08, BST01, BST03, BL92, BFGM05, Ban98, BP04a, BHH04b, BHH04c, BBH^{+}05, BGH06a, BF08, BTQ02, BD12, BM11a, BC92, BJP12, BPR02, BL08, BHRU09, BCG^{+}07, BB01a, Ben06, BP09a, BV92, BR89, Ber98, Ber02b, Ber00, BDH^{+}99, BCP99, BCI85, BHCL87, BM99b, BK80, BC93, BD97b, BG12b, BSBM05, BTBP^{+}08, BTPPS09, BJN03, BKLP11, BL89, BCC^{+}12, BTF02, BF89, BL02a, BM09a, BM94, BMYZ91, BMYZ95, BMCCM11]. France [BCRT92, BFMR96a, BFMR96b, BFT^{+}08, BBMSD09, BDN08, BP09b, BKM01, BGKS04, Bra99, BLV02, BKZ05, BB02, BV05b, B^{+}12b, BBB06, Bun94, BDLMM12, BH06d, BFK^{+}08, Cal82, Cal86, CBC^{+}02, CDV03, CHD07, CP12, CS03, CP10a, Cas01, CJ08a, CD07, CL07, CGP02, CT07, CMZ99, CM03a, Cha09, Che99a, CHG^{+}07, Cho88, CL90, CCBW11, CHAL04, CCLR98, CLB11, CCM08, CSTD08, CG88, CW89, Coh94, CGM95, Coh96, CMS11, Coi99, Col02, CML^{+}10, CJ95, CBF01, CMT01, CP07, CW88, CFP94, CSRS08, CCR86, Cou01, Cou05, CM11, CvRVSN12, CDS09, CDT11, CBCGA12, Cup00, CD11, COB^{+}08, DL06a, Dam99b, DS05a, DSHL11, DN88, DH97, DMR01, DFPV11, DE05, DMW00, DRDKE11, DPV11, DL04a, DJ90, Des86, DH10a, DEQ92, DKLT05, DEM96, DOS99, DH04, DC00]. France [Dit01, DGG^{+}10, DTF98, DLM07, DLJ12, DT06b, DFGN12, EFG^{+}09, EH08, ECB06, ELNM08, EMW94, Est99, ES92, FRS08, FL00, FLS06, Fau90, FD03, FKP99, FR95b, FZJ01, FPMW06, FJ92, FLP12a, FO95, FvdTDG12, Fon00, FTZ08a, FTZ08b, FTZ08c, FTZ08d, FHP03, FS07d, FW09, FZ86, Fri01, FLP11, FCDB07, Gab13, GP05, GP07, GP09a, GP11, GPAS^{+}06, Gal97b, GLP12, GG05b, Gan88, GANACBR10, GAL12, GG09, GS01, GLV09, GL10a, GJ93, GNR07, Gau93, GPR^{+}12, GM12, GCK06, GH05, GPM98, GLM^{+}08, GOZ05, GZR02, GP89, GO06, GP09c, Gue90, GQ95, HRZK02, HS12b, HT97, HCT02, HLSZ11a, HLSZ11b, HT11, HMT10b, Han07, HLR^{+}98, HLC^{+}12, HKM95, HR95, HIS05, HAB98, Heu98, Heu00, Hig05, HW03, Hor00, HCDL09]. France [Hor02, HSMPA09, HKS09, HM98, HDLM09, IHQ96, JNR11a, JNR11b, JP09b, JVA04, Jou85a, Jou85b, JK06, JRR03, Kah79, KMP84, KWA02, KJ88, KS02c, KLM^{+}07, Ker97, KBP07, KW90, KR00a, KR02b, KJB09, Kla12, KW98, KNP01, KS07a, KK09b, Kot86, KT11, Kra05, KB92, KS91, KU09, KSGT10, KR04b, KS07c, Lac10, LV12, LY04, LSW01, LN04a, Lau93b, Lau00, Le 02, LLP99, Les87, LPZ12, LLP^{+}07, LCF^{+}04, LS89, Lor01a, Lor01b, LD05b, Mag03, MMM01, Mal97a, Mal02, ML04, MR97a, MVL11, MT07a, MT06a, MT06b, MTH06a, MTH06b, MM10, MD96, MKZB04, MMU96, MS12c, MKB^{+}09, MVN86, MTC^{+}08, MR95, Mon96, Moo01, MM97, MMK98, MCdS^{+}08, MC98, MPS04, MSR99, MNW98, NP02, Nad99, NNS03, NPPDG06, NP09, ND09, NV08, Ng79, NP10, NS87]. France [NE02, Nie09, NL99, NvRSS11, eOP05, OO12, OLL10, OS04, Oli07, Omi00, PR08, PAD^{+}08a, PV00, Păs09, PH10, PMA^{+}07, PR05, PT08b, Pin89, Pit91, Pod95, PP11, Pol86, PMS08, PR96, PPF^{+}00, Ral94, Rao92, RT00, Ret97, RMTW11, RR04b, Roc07, RH01, RBC93, RS03, RS01, SRG04, Sd98, SVS02, Sch10a, ST08, ST11, SDR^{+}00, Shv06, SCG98, Sif90, SA97, SDH03, SRT06, SP12, Sub05, SC07, TLC06, Tau08, Tho06, TC98, TND10, TL02, TST09, TGM97, Vää99, Van97a, VMvN12, Vau98, VA07, VZM^{+}10, VAM06, Wan98, War94a, WK02, WHG07, Wii04, WJP07, WFB12, WWD99a, WWD99b, WZ11, YBM87, YLS^{+}12, YHV12, Zam08, ZKZ00, ZQ98, dBBM09, dGEK11, dCHM12, vLS11, vdABCC05, BGY96]. France-Japan [BGY96]. Frances [MP93]. Francez [GKKW09]. Francicsco [MS02c]. Francisco [AD05, Abe06, ADJ^{+}01a, ADJ^{+}01b, DS03, Dun12, Fis09, GAA06, God05a, GZ09, HH03, HW08b, IR08, Joy03, KS12b, Kia11, KMSS02, KP15, LSW02, LSgL^{+}06, LPZ08, LHH11, MWYF10, MT09c, Mal08, Man75, MTO99, MRY^{+}11, Men05, Mue01, Nac01, Oka04, Pie10, RLL04, Rei09, SWB03]. Franco [DEM96]. Franco-Chinese [DEM96]. Franco-Japanese [DEM96]. Frankfurt [BH00, CE95, DZ10a, GSW06, LLMT04, MF11c, PPCS12, Tou96]. Frankfurt/Main [CE95, DZ10a, GSW06, MF11c, Tou96]. Frascati [GG02a, Paz97]. Frauenchiemsee [FKKB^{+}05, WPH03]. Frauenwörth [BMS08b, BFKO05]. Fredrikstad [WGD05]. free [Nij80]. Freiburg [BHN97, BN98a, BN98b, CHZ08, FNGW08, KvG98, Mue06, Nic86, dRF01, dRS01]. French [Coh92, Coh94, CJ95, CKOS09, Gil10, PH04, TTF07, Wri03, GB77]. French-Israeli [Coh94]. French-Japanese [CKOS09]. French-Soviet [Coh92]. Frequency [Yal10]. FRG [BB89, BHL90, Ham91, Bla90, BHR90, Bod91, BBR88, BBR89, BBR90, Dit88, FM91, GS88, Her91, Kar91, KP92, MT86, MC89, Neh88, NS90, Rei89, SPZ90, SH91, SM91, Sti90, Stu90, Vog88]. Fribourg [FK95b]. Friedrichshafen [KW96]. Friendly [LBM^{+}10, MSRP98]. Friuli [MSH08]. FROCOS [Gra05, GS09a, KW07c, TSS11, Arm02, KR00a]. Frontier [Sex09]. Frontieres [Gau93]. Frontiers [CHB^{+}12, CWiM^{+}09, GCSR07, HKI^{+}09, NMM10, NBGA08, OBS12, OBM11, SHN07, SINK08, TNN^{+}01, TNA^{+}10, WSN^{+}06, WSTI07, ZYW08, vLWH^{+}00, Arm02, ALZ11, DHX09, GS09a, Gra05, KR00a, KW07c, LCY10, M^{+}06, PF07a, PWY08, SLSW12, THX^{+}07, TSS11, ZLS^{+}06]. FRSM [YAKM^{+}12]. Frundlagenforschung [BHR84]. FSE [Bir07, Can12, Dun09, GH05, HI10, Jou11, NL99, Nyb08, Rob06, Bih97, DR02a, JS97, Joh03, Mat02, RM04b, Sch01a, Vau98]. FSE’99 [Knu99a]. FSEN [AS07a, AS10, AS12]. FSKD [WJ05, WJS^{+}06]. FSMNLP [YJKK06, YJKSW10]. FST [AS02a, HMV01, KP00, PR03a]. FSTTCS [AKG06, AP07, LM04, RS06]. Ft [OPS11]. FTRTFT [LY04, DO02, Jos00, RR98a]. Fuji [HW06a, MS99b]. Fuji-Susono [HW06a]. Fujian [X^{+}12a, X^{+}12b, XBM12]. Fujiyoshida [IRM93]. Fukuoka [AT02b, AM98, HY12, Pol97, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, UWB^{+}06, XSA^{+}12a, XSA^{+}12b]. fulfilling [Luk97]. FUN [BG10b, CPP07, KKL12, BG10b, CPP07, KKL12, MdRIR08]. Funchal [JJF03, MPF95]. Functional [BAL88, BGH06b, CHZ08, EKLM87, GH12a, HM11b, Hor06, HZB07, HPSZ08, HPZ10, KPS09, MS11a, PHZ11, PP12b, Sch03b, SC11, ZHP12, AM02, CHD98, CCR86, GHMT05, HDC99, JM95, JJ03, Klu97, KC00, LMS96, MK01, PA03a, SHO99, TMP04, VU05, ADS09, BKV10, Esc10, FRSH05, GH08, HW06a, HP95, HRA02, Hug91, Jou85a, Kah87, KS04a, KMC^{+}01, KU01, Kuc11, Lau93a, Mag03, Mar11c, MA11, MS99b, SS07a, ST12a]. functions [Tha84, Bol99, Dav81, ST87]. Fundamental [BH06b, But91, Cer05, CW09, DL07, Est98, FI08, Fin99, GO11, Hus01, Mai00, Pez03, RT10, WK02, WMS04, dLZ12]. Fundamentals [AS07a, Bis88, DT89, HKV10, JMC10, LN03b, Mit05, Mül07a, TDG91, Gog94, AS10, AS12, BJ86b, Bud85, BBL87, Bud91, CC97a, CP99, CDG89, CVÉ07, Esi93, Fre01, Gec81, Kar77, KJ91, KCG09, LR05a, OST11, Rei95]. funfte [Müh75]. Fusing [Har11b]. Fusion [DKN^{+}11, LSR03]. Future [ÁCD^{+}12, BRS^{+}07, BGPTF10, CCNY10, CS10, CRS00, DFT09, DGG^{+}11b, HTMC09, HHG11, KS03b, KD11, LL06, MG07b, OS09, PPPM10, SF11, Smi03, SPSV^{+}04, SSKR02, VS09, Whi00, WKB04, XCR11, ZSP10, BFS92, Che99b, II00, KK00, Pla96, Pup99, RBS93, Smi01, SSv94, BV92, BGPTF10, DFT09, hKhLKŚ10, hKAS^{+}11, hLhKcFŚ09, MZCC09, Sch03a, TV87, ZCCL10, ZSP10]. Fuzzy [ASR^{+}07, BdBK03, KŚHM11, MMP07, MCA^{+}07, PS02a, Reu01, Rey04, SCH^{+}09a, SWS^{+}05, SYP^{+}05, CKPB07, DER94, Fur95, PSD^{+}05, Reu97, Reu99, Wan03, BdBK03, BPT06, DPP09, DER94, FPP11, FU96, GMP06, KS93, Ral94, RM97a, RS99a, WJ05, WJS^{+}06].

GA [ASE^{+}07, BBD^{+}10, CDA^{+}06, FS10, FHP11, GS09b, JMO09, KMR^{+}05, MZ07b, MSZ08, SR11, SvASD05a, SvASD05b, SvASD05c, vdMS11, HNM02, Pic01]. GADA [MTH05, MT06a, MT06b, MT07b, MT07c, MT08a, MT08b]. GAG [KHZ82]. Gainesville [HW06b]. GAIW [NSM07]. Galdana [AS06]. Galinhas [BR02, FLS05]. Galinhas/Recife [BR02]. Gallen [WZA10]. Galway [AST05, GMBM05, PdA^{+}11, RG11, tTVH^{+}12]. Game [ABB10b, BKA11, BEJvR11, BLvdH10, CKK^{+}09, GW12, KKS10, MP09, MS08b, Per11c, Ser12].

Game-based [CKK^{+}09]. GameDays [GMUW12]. Games [AF11, BBDS12, BEJvR11, BCK10, CSCR08, CKK^{+}09, CHCM11, DBSvD09, Dig11, EKO08, EGO09, GMUW12, GTW02, HBN06, Kai11, MOP11, MOH^{+}12, MdRIR08, MF01, SMB03, ZZP^{+}10, vdHHsHD06, vdHCD07, vdHXMW08, vdHS10, vdHIP11, vdHP12, vdHI99, vEV12]. GameSec [ABB10b, BKA11, GW12]. Gammarth [Pre09]. Gan [HBN06]. Gandhinagar [JM11b]. Gap [MHS10, PST11]. Gaps [HKP10]. Garda [BVAT12, EL07, FdS12, NVSB^{+}07]. Gardens [HNM02]. Gargnano [BVAT12]. Garmisch [EB04, SP93]. Garmisch-Partenkirchen [EB04, SP93]. gate [GH93]. Gaussig [GLW88]. Gaze [EV09]. GB [MJL02]. GbRPR [BV05b, EV07, JFT11, TEB09, HV03]. GC [PQ05, Pri03b]. GCB [Hof97]. GCC [LSDN04a, LSDN04b, ZF05, JPXS04a, JPXS04b]. GCCB [DSS^{+}07]. GCIC [CPGL05]. GCSE [Bos01, BJ01d]. GCSE’99 [CE00]. GD [BC11c, EG10, HN06, HNQ08, KW07a, Pac05, TT95, TP09, vKS12, Bra96, GK02, Lio04, Mar01, Nor97]. GD’97 [Di 97]. GD’98 [Whi98]. GD’99 [Kra99]. Gdansk [BS93, Gór06, JS05, Maz76, RFBW12, WDKW08, KY11]. GDB [XYZU11]. GDC [hKAS^{+}11]. GDM [YMY^{+}10]. GDMM [BGG^{+}12]. GDR [AJM87, BJ86a, Bis88, Dav89, GLW88, Ham89, Jan89, WLS89]. Gdynia [JNHJ10a, JNHJ10b, JNH11a, JNH11b]. GECCO [CPFD^{+}03a, CPFD^{+}03b, DPB^{+}04b, DPB^{+}04a]. GECON [ANF08, ABR09, AR10, VAR12, VA07]. Gedi [BtHRBB12, ER12]. Gelfond [BS11b]. Geltrú [CPS05, MBVTY11]. Gemany [May01]. General [Bra80, STL11, Som09, Pfi96, ABC^{+}06]. Generalized [Vää99]. Generating [KvETvdH10b, KvETvdH10a]. Generation [ASE^{+}07, ABK12, AH07, BMK08, BMK09, BDK10a, BKH11, BEP04, CA08, aCF06, CW07, DPLW08, EKM06, FKK09, GVCA07, HG02, hKhLKŚ10, hKAS^{+}11, Kon04, KHI06, KHS07, Kou11, KT10, hLhKcFŚ09, LZ02, MCA08, OSC^{+}08, RS00, RST09, RR04b, SSN08, Tah00, Ull08, VC03, AZ96, Dal92, Jon80, KS05b, LBCO04, MSY00, Pai00, PZ92, PT99, SS91, Tah01, CHCA09]. generative [Ley01, BCT02b, Bos01, BJ01d, CE00, FLVS11, GL05, KV04, LSV06, LVS08, PS03]. generator [KHZ82]. Generic [JLM00, Pau94, WSL93, BG03a, BGHJ07, Gib12, JLM00, Mel04]. Genesis [YKMJ09]. Genetic [CPFD^{+}03a, CPFD^{+}03b, DPB^{+}04a, DPB^{+}04b, STWS07, WVDS05, Fur95, Ban98, CPFD^{+}03a, CPFD^{+}03b, CTE^{+}06, DPB^{+}04b, EOE^{+}07, EAES^{+}10, Fos02, KOL^{+}04, KTC^{+}05, Mil01, MSK^{+}12, OVG^{+}08, Pol99a, Pol00, RSK03, SFN^{+}11, VGM^{+}09]. Geneva [AK07, AP09, Ame91a, BCT02a, FR97, GB07, KLPP03, MTT98, MMBND07, NJ10, TWN92, Tro11, WMR00]. Genoa [AB81, DP11, Dro09, FSA^{+}07, Küh07, MPT10]. Genomics [BEM06, CM09b, EIRW06, Lag05, NV08, Tan11, TD07, EW05, MH05]. Genova [CV04a, CR01a, HM01, Hus01, MS03a, MY01, NWHR06, San01, Wil01a]. Gent [BD03, DW12]. Geographic [Bar02, KWT03, FCF92, FM99, Mon01, Nie94, Ooi90, Sch97b, Vck99, van97b, CMB^{+}08, EM02, EFM04, FRvKS10, RMFG06, XKGS12]. Geographical [BRL08, CT06, CFM09, DPT12, TFK11, WT07, KBC05, LV05]. Geometric [CJ08b, LOS05, MSX10, Gun88b, LM96, PFCS05, KS06, Lau93b]. geometrical [CRYS09]. Geometry [AK03, AKT06, ACK^{+}07, ABKT11, BIK01, BR07a, BRP09, CSTD08, DRDKE11, Heu04, IKKU08, KUA00, KMS12, KNP06, MRU12, NSS03, SNRG11, SZ11, Win04, AF97, AKU01, ABK05, ADL05, BCP99, BN91a, BNS00, BLV02, GWY99, Hel91, Her92a, HW06b, LM96, MMU96, Nol88, RGW01, Sch97b, Wan98, AKR03, BL89, Tei93]. Georgia [ABS01, BV02, BLSS11, BGL09, GVM99, WGL96, YD98, tCZ07]. GEOS [RCLE05, CLB11, FRL07, JRL09]. GeoSensor [NLS08, TMN09]. GeoSpatial [CLB11, FRL07, JRL09, RCLE05]. German [BBE01, BE11, BLKP08, BFP04, BI73, BvdHPP09, BHN97, BHH74, BCC99, DBB^{+}08a, DBHS10, DW10, EDD04, EKL^{+}05, FTAZ06, FKS07, Fur05, GK12, GH96, GKN03, HS74, HBE07, HG98, Hof97, JKL02, KO11, KF74, Lei74, LDTU04, LSU07, MHA09, Müh75, NDF94, Ohl93, PMKG07, Pup99, RBW95, SKMT04, SF74, Sch74, Sie75, SS76, TG12]. German/Austrian [BBE01]. Germany [ACdR^{+}08, AL11a, AKK04, ABMM07, AB11, ASS10, AMU03, AV07, ABB10b, AH03, AFKL04, ANR99, ABB99, AFM01, ADB^{+}05, ABMH08, AJB09, ASLM09, Amb01, ABH^{+}10, ADMH04, ADM^{+}06, ADM^{+}08, AAB^{+}07, AJ94, AL12, AWA^{+}11a, AWA^{+}11b, AKR03, ABW10, AH10, BE11, BH06a, BdBC01, BS09, BCD^{+}03, BKLM07, Bar11, BBAB^{+}07, BRS^{+}07, BPR05, BGV01, BMP01, Bau99, BHP95, BE87, BEM91, BPR08, BWAM09, Bec05, BHGP02, BCRB^{+}11, BB95, Bel03, BR92, BS03a, BHMS07, BS92, BMS04a, BMS08b, BNUW09, BLKP08, BGPTF10, BF11, BLB^{+}03, BGD^{+}05, Ber03, Bes93, Bet83, BFS92, BG11b, Bil97, BMS95, Bis02, BH92, BL07a, BFP04, BH77, BZ83, BRW93, Bod96, BLKW01, BU01, BHU03, BM02a, BJ01a, BKS01, Bor91, BP05, Bos01, BHJ^{+}11, BE07, BRM06, BvLR04, BFH98, BTT12, BVNW87]. Germany [Bra96, BW00, BC11c, BL01, BKM07, BvdHPP09, BH00, BJS93, BHN97, BHLvdP07, BUHS08, BR98, BS12, BK04b, BSL^{+}09, Bud91, BLPW02, Bun96, BCC99, BE01, BLVV07, BN98a, BN98b, BL06d, BBC04, BRS09, BJ01d, BMP96, BGW06, BKO03, BFKO05, CC93, CL96, CL08a, CE95, Car03, CZ02, CHPW05, Ceb08, CBNR02, Cha03a, CSR08, CHZ08, CBN99, CS00b, CS01a, CC09b, CS96, Cov99, CL08b, CBBS09, CMOZM11, CGB11, CRS00, CKLM09, CE00, DO02, DH07a, DGT96, DEKR09, DG01, DZ10a, DGG^{+}11a, DFMV10, DFP^{+}12, DKM91, DH11, DBB^{+}08a, DNS09, DPW04, DLN10, DVM^{+}11, Die02, DD05, DN09, DMWW88, DBHS10, Din03, DJS91, DFN97, DPP97, DdCF98, DFN10, DW10, DRS05, DHMM96, DW07b, DHH^{+}08, DA01, DBB^{+}08c, DHSH96, DS07, EV96, EJM10, EHP94, EB04]. Germany [Ehr98, ENR83, EKR91, Ehr91, EEKR00, EEKR12, ES02, EFW93, EBT07, ELS09, ENO12, Ern07, EKL^{+}05, FS05, FS99a, FR09, FR01b, FMW01, FS99b, FTAZ06, FDL10, FBM12, FST09, FLUV02, FMNS06, FSTS09, Fre98, FM99, FKKB^{+}05, FKS07, FNGW08, FLR11, FW92, FLMS05, FLMK06, FLT07, FKLT08, Fum97, Fur05, FSS06a, FSS06b, GO94, GO96, Gab97, GM00, GMV07, GL97, Gel99, GWS05, GH94a, GH94b, GKS^{+}07, GKMV11, GK06, GL10b, GKT09, GKS12, GHP^{+}11, GBC^{+}09, GO11, GKL^{+}11, GK12, GSH^{+}04, GMI06, GMUW12, GRK^{+}10, GMS06, GLIC10, GF12a, GH96, GR06b, GBMS11, GS00b, GSW06, GHMT05, GT05, GSS00, GHI^{+}06, GNR94, Gue01, GKN03, Hac75, HP12, Hag02, HM08, HSJ07, HWP01, HR96, HG96, HK05a, HS76, HS08b, HU08, HLW04, HRB12, HMM12, Her92b]. Germany [HHJZ12, HHLR08, HBE07, HR91, HG98, HPP06, HSS07, HR08, HW07, Hof97, Hof11, HK96, HLW08, HKP11, HMMW10, HNW04, Hsi95, HKH10, HLFS12, HS98d, HSTU99, HMSU04, HU05, IPSK04, ISS^{+}06, JMBS07, JK00, Jan92, JLST06, JO99, JKL02, JLM00, JZ03, JKB04, JLS11a, JLS11b, JCC^{+}12, JHC02, JP09a, JFT11, Jou94, Jou09, JK05b, KSHS01, KMTW03, KWA01, KK97, KWW04, KN91, Kat99, KK11a, KS09b, KW07a, KGRD10, KIRK05, KEM07, KSW05, KM04, KK98, KHG01, Klu97, KO11, KOK04, Kno11, KvG98, KFW04, Koh06a, KK09a, KHW03, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KW10, KS05b, KGS08, KMTU03, KP05, KJ10, KS96, KMN94, KK95, KW96, LPT07, LM98a, LSS02, LLMT04, LdRV94, LK99, LS04a, LMSS07, Leh11, Len93, LGG97, LBCO04, LS05]. Germany [LM08a, LW09, LDTU04, LPH00, LR05a, LK07, LK09b, LSW08, LS12b, LKK10, MOH^{+}12, Mai00, MP06, MSG07, MG08, MRK05, MNS05, MW91, MS96a, MVC07, MW11, MR12a, MT97, May93, MST95, MP95a, MT99a, MHA09, MF11c, MMR93, MM96, MK03a, MS06b, MP12, MK01, MTWD06, Moh91, Moh97, Mol95, MF02, MS08b, MBS05b, MK12, Mue06, MCER10, MSUB04, MSKY10, MdSC09, MYS10, NWL06, Nag95, NW01, Nar01b, NDF94, NR98, NKPH08, Nic75, Nic86, Nie93, ND04, Ohl93, OZ01, OTZ01, OD12a, OD12b, Ott87, ÖRH12, PRV11, PPCS12, POP09, Pat00, PW92a, PW92b, PpWR96, PP99, Per01, PBH03, PR03b, Per05, PI05, Per06, Per07a, PS07a, Per07b, Per08, PS08a, Per09a, Per09b, Per10a, Per12a, Per12b, PT92, PBGK02, PMKG07, PS03]. Germany [Pfe07, Pfi00, PRdRW09, PRG09, PK00, PS11b, PW07, Pup99, RF01, Rah04, RBGS04, RMFG06, RLKHL12, RR03, RBG09, RD12a, Rei95, RST09, RM97b, REF^{+}06, Reu97, Reu99, Reu01, RMS^{+}05, RSS06, RB91, Ric98, Rig08, RABR03, RW86, RBW95, RBS93, RHD08, RB10, RPZ97, RH98, RFBD09, RJL^{+}08, RB00, RME04, SST00, SO07, STP^{+}10, ST00, SH08, SDG^{+}07, SKP07, Sch93a, SKMT04, SH12b, SSW92, SUW02, Sch12, SE94, SB07, SV97, SNRG11, Sch92, SSN08, SNZ08, SM06, ST12b, SHGB05, SHW09, Sen10, SSD^{+}08, Smi05, Smo00, SOP12, SK97, SDP97, SZ00, SP93, Som96, SEN07, SPB00, SLPM96, SLKL09, SS08b, SSU12, SHL08, Ste95, Ste01, Ste94, SW97, SP07, SRM^{+}11, SCKR03, SLP05, SFV^{+}11, SKB98, SDM05, SH11, TY11, TDG91]. Germany [Tha96, TW07, TG12, TSS11, TS87, Tiu00, Tou96, UBM02, Vah09, VGM^{+}09, VDD^{+}11, WF98, WK08, WRP07, WKP03, WPR08, Wat00, WL04, WAR94b, Wet92, WBD08, WP10, WSG07, WN96, WPH03, WHR10, WHS01, WPZA11, Wol07, ZK06, Zan00, ZJ04b, ZT04, ZSP10, dMB05, dMS06, dMLG03, dRF01, dRS01, dRLP98, vO04, vM94, von96]. gestalt [Lem97]. Gesture [Bra99, CV04a, DGWB09, EKF12, GCK06, KW10, WF98, WS02, Bra99, CV04a, DGWB09, EKF12, GCK06, KW10, WF98, WS02]. Gesture-Based [CV04a, DGWB09, Bra99]. Gestures [EV09, ZZTG07, ZGT05]. Gestützter [BHH74]. Getting [EL02a]. GEVO [JPXS04b]. Ghent [BTKP^{+}11, CW11, CDY^{+}11, DKS^{+}07, DLNU11, DB04, DY10, TWSN07]. Gheorghe [KK11c]. GI [CK82, Deu81, KEM07, Lei74, Müh75, MCER10, Sch12, Sie75, Wei79, BB95, Bra75, LM98a, Lei74, LPH00, Müh75, Sie75]. GI-4 [Sie75]. GI-4. [Sie75]. GI-conference [CK82, Deu81]. GI-Dagstuhl [KEM07]. GI-NTG [Lei74, Lei74]. GI/ITG [MCER10, Sch12, BB95]. GIAE [LOS05]. Gießen [HKP11]. Gießen/Limburg [HKP11]. Gif [Bra99, CMT01]. Gif-sur-Yvette [Bra99, CMT01]. Gilles [KR02b]. Girona [AK12, KCMO11, MBMS07a, MBMS07b, MOFM10]. GIRPR [MST12]. GIS [AS99, FC93, FK95a, HF97]. GIScience [CMB^{+}08, EFM04, FRvKS10, RMFG06, XKGS12, EM02]. Giss [JPXS04b]. Gjøvik [SG03]. GKR [CRW^{+}12]. Glasgow [Bla05, BMS01, BD04, CK07, CGvR02, CR05b, CFS06, DJNR06, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, Joh01, LJR^{+}10, LIH99, MOP^{+}08, MSH85, MB06, Myc95]. Glenwood [CS05b]. Gliwice [PK12]. Global [AKTT12, Ayk07a, Ayk09, BF08, BJN03, BS12, CKH10, DKW07, KN09a, KT11, MSB07, Rau11, Ria08, SZC^{+}03, WHR10, DS05b, FP90, PR87, BJN03, JNS05b, Pri03b, PQ05, TZ89, WZA10]. Globalized [CC12]. Globally [WPR08]. Globe [Ham08, HT09, HMT10a, HT11, HHMT12]. GLOW [AKTT12, KT11]. Glyfada [Sel95]. GMP [CJ08b, KS06, MSX10]. Gmunden [Mio93, Vol93, vdV84]. Goa [APBP07, BPSP06, DBM09]. Goal [Fri06]. Goal-Directed [Fri06]. Goar [BHP95]. Goast [KLC07]. Godel [GLM93, GLM97]. Godesberg [Klu97]. Goedel [BM03a]. Goguen [FJM06]. Going [BL10b, GZR02]. Gold [Boy01, CKH10, HNF08, KLC07, OT07, Pap95, RAW07, SZ93, WLP^{+}09]. Goldwasser [Gol11b]. Golumbic [LLM09]. Gosen [Bud91]. GOSLER [JL95]. GoTAL [NR08]. Göteborg [BH02, PVC^{+}99, Pol99a, SDK^{+}08]. Gotemba [LMMN10]. Gothenburg [AS07c, Gud08, LK10, NR08]. Göttingen [SPB00]. Governance [KLLP12]. Government [AFGvE10, AFGvE11, JSWhT11, KLLP12, SJW^{+}12, Tra04a, Tra04b, Wim03, Wim04, WSGA06, WSG07, WSF08, WSJT09, WCJS10, TL02, WTGA05, BGPW05]. Goyang [KBC05]. GPC [AP09, BCC^{+}10, CL07, CM06, LCB12, RKM^{+}12, RYG11, WYX08]. GPCE [GL05, BCT02b, KV04, PS03]. GPS [FK09]. GPS-less [FK09]. GPWW [ED06a]. GRACE [M^{+}06]. Grading [Sta99]. grain [Boc95]. Gramado [HP09, LCD^{+}09, OW09]. Grammar [PQR11, Woo80, dGEK11, dGN12, Hau86, Ste12, Uhl82, Yel88]. Grammars [CEER96, Kuh10, MR12b, AM91c, DJL88, DJ90, EKR91, Hab92, Nij80, Ste12, CER79, ENR83, Ehr87]. Grammatical [AFvZ02, CO94, CCM08, HS98b, MD96, Oli00, PS04a, SKS^{+}06, SG10]. Gran [AAB^{+}06, ANF08, CAS^{+}09b, CH99, DPA05, DPA07, LMB08, MDP04, MDBF01, MDPQA09, MDPQA12a, MDPQA12b, PMD97, TGAH11, VSH11, vdL98, vdL01]. Granada [CKM93, LSW03, MAB^{+}02, MP01a, MP01b, Mor95, Pri91, PP04]. Grand [CDCW09, Syv02]. Grande [CDGW10, CFVV^{+}12, CEM^{+}06, FdCLJ^{+}11, dSK12]. Granular [ASR^{+}07, KŚHM11, SCH^{+}09a, SWS^{+}05, SYP^{+}05, Wan03, ZSO99]. granular-soft [ZSO99]. Graph [ACK^{+}07, BJ12a, Bra96, FHP11, IKKU08, KS10b, Leo04, Nor97, PNB04, SNZ08, A^{+}08, ABK05, ADL09, BC07, Dor95, EEKR00, Hof82, NSM00, Pau93, SVV12, BS92, Bod03, Bra96, BW00, BC11c, BL01, BKM07, BEFP08, BV05b, CER79, CEKR02, CEM^{+}06, CRW^{+}12, CEER96, Di 97, ENR83, Ehr87, EKR91, Ehr04, EHRT08, ERRS10, EEKR12, ELS^{+}10, EG10, EV07, FK87, Fom06, GSLM12, GK02, GS88, HV03, HN06, HNQ08, HS98c, HNW04, JFT11, KW07a, KK11d, Kra99, Kra05, Kuc̆02, Lio04, LLM09, Mar01, May93, MST95, Moh91, Moh97, MJL02, Nag90, Nag95, NDFMS97, NS81, Nor97, Pac05, PH10, SE94, TT95, Thi10, TS87, TP09, TEB09, Whi98, WNE99, vKS12, vL89, vL94b]. Graph-Based [EV07, JFT11, BV05b, TEB09]. Graph-grammars [CER79, ENR83, Ehr87]. Graph-Theoretic [Bod03, BW00, BKM07, BEFP08, Fom06, GSLM12, HNW04, KK11d, Kuc̆02, PH10, BS92, BL01, GS88, Kra05, May93, MST95, Moh91, Moh97, Nag90, Nag95, NDFMS97, TS87, WNE99, vL89, vL94b]. Graphics [BKW09b, BTCW10a, BTCW10b, BTCW12, BKO04, BFKO06, BFK^{+}07, BFK^{+}08, BFC^{+}09, DVM^{+}11, GP07, GP09a, GP11, KP06, NPS06, TBKO10, BKO03, BFKO05, Chi95, OK90, BK02, CD00b, KT96, LL06, LLO08, LK04, OLL10, TC98, Her92a].

Graphs

[ABKT11, Dau03, DDM08, KW01, MMS93]. Graphtheoretic [Nol81]. Grassau [Rei89]. Graz [AdBB12, BS06b, BW97, CKM^{+}09, CBW99, Cro03, FJL97, Hic04, Hol07, Hol08, HS11, HSL06, LBP06a, LBP06b, LBP06c, LBP06d, Mau79, Mau91, PPBL12, Rob06, SKMR94, VHM07]. Great [MTNL12, SdDK91]. GREC [LL06, LLO08, OLL10, BK02, LK04, TC98]. GREC’99 [CD00b]. Greece [AK97, ABK^{+}09, AB12, APS08, AMSM^{+}09a, AMSM^{+}09b, APS06, AGS^{+}11b, AGS^{+}11a, AZ11, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, AOSV08, ACHH12, AKSO06, BBP^{+}12a, BBP^{+}12b, BDL08, BDW08, BDSW09, BCP^{+}04, BH05, BR07b, BR09, Bra85, BCKT05, BDFS04, CL11, CV08, CCPP12, Chr91, Chr00b, CCK95, CMP07, CMV96, CD01, CDND11, Cou93a, CLBP11, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, DVVA08, DKK08, Del00, DDI10a, DDI10b, DDI10c, Dol09, DFM11, DKW07, EKF12, EKP^{+}07, EHH12, EM06, EBAH08, Gae96, GMV05, GANACL11, GCFA12, GVT08, GPE05, GPT10, GS06, GHMV11a, GHMV11b, GHMV11c, Had03, Hal94, HPTV05, HPP88, KPS12, Kar94, KGL04, KLB^{+}06, KSDO06b, KSDO06a, KPK^{+}10, KKS10, KMMT12, Kut98, Kut12a, Kut12b, KA05]. Greece [LW95, LM06b, LMT^{+}04, Lop06, LCDT08, LDC^{+}09, MPK04, MCKB06, MPV12, MMMP12, Mak86, MPS06, MSS12, MS10b, MS10c, MS12b, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, MKR^{+}04, NT06, Nie99, Nik05, NCJK08, NT03, ONPR07, OSvL01, PT09, PS04a, PGA^{+}05, PR11, PAD08b, PW09, PDK^{+}10, PCG^{+}10, PV04, Rei88, RG03, Ria09, SH12a, SMBQ07, Sel95, SCP^{+}12, SB12a, Spi95, SMK06, SS06d, SZ05, SBH^{+}04, Thi10, TP09, TNP^{+}07, TSK92, TZS00, VWH06, VBH07, VEG^{+}11, VS02, VP04, Wim03, XW11, vdMBD08]. Greenbelt [HRT^{+}03, HRTR05, HRR^{+}06, Ras01]. Grenoble [ABFV12, Alo05, BP09a, Ber02b, BL02a, BM09a, Cal86, DE05, DPV11, Hig05, Hor02, KS02c, Ker97, LY04, Lau93b, Le 02, Mal97a, Moo01, NE02, Păs09, PR96, RR04b, SVS02, Sif90, TGM97, WK02, vLS11]. GRID [BB00b, AP09, BCC^{+}10, BGKS04, BSW12, CL07, CM06, DSS^{+}07, Ham08, HT09, HMT10a, HT11, HKW11a, HHMT12, KS05a, LCB12, RYG11, SF10, SHP^{+}05, WYX08, HPR05, JM02, TAS05, ANF08, ABR09, BB00b, Dik04, JKB04, JPXS04a, JPXS04b, KS05a, Lee01, LSDN04a, LSDN04b, Par02, PF07b, RKM^{+}12, RBTD04, SF10, VA07, ZF05]. GridGIS [M^{+}06]. Grids [AR10, RBTD04, VAR12, Pie05]. Grimstad [PRR05]. Groningen [BIMW12, PW03, Van89, WR09]. Gros [Dan12]. Grounding [VSTN06]. Group [AD98, BS10, CHD07, CD01, CDND11, CH01, DKD05, DLM99, DKP00, KGRD10, KKDV02, KKD04, LdRV94, LKD08, MTWD06, MT99c, RWD09, SCKR03, SBH^{+}04, TBD12, WK08, BDW97, Bui87, DLO03, RF99a, Hof82, Len90, SCKR03]. Group-theoretic [Hof82]. grouping [FMdGC99]. groups [But91]. Groupware [BAdVR08, CBF09, DZGS06, FD03, FLS05, HP02a, HOC07, dVGR04]. Grove [GHR89, SWAF95]. Grzegorz [BEKS02]. GSEM [JKB04]. GSN [NLS08, TMN09]. GTP [M^{+}06]. GTTSE [FLVS11, LSV06, LVS08]. Guadalajara [BCE09, BRU06a, RUL04, RRU05]. Guadeloupe [TTF07, Wri03]. Guanajuato [ABG09a, CAZ05]. Guangzhou [BW11, BZ08b, CFK^{+}12, LLT^{+}04, PZ09, SKK08, WmCL07, XZCL07, YLWY06, vDLJ11]. Guarda [Fil91]. Guarded [HR99b]. Guide [BHH^{+}04a, GTW02, ALM95, GBDM77, Hou90, SBD^{+}76, Smi74]. Guided [MDHB11]. Guidelines [AH97b, RtTMP10].

Guildford [BMPM09, Bow94, HSKB09, KN03, Pit86, TKHS05, WR03]. Guilin [HhKRT06, LLW06, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, SS06b, Wan01, YW06]. Guimarães [BL12, IZP09, NSM07]. Guixols [PB97]. Guiyang [LA05]. Gullane [ST88]. GULP [DP10b]. Günter [HHG11]. Günzburg [JMBS07]. Gurevich [BDR10]. Guwahati [CDPT06, VG12]. GW [DGWB09, EKF12, GCK06, KW10, Bra99, CV04a, WS02]. GWAI [Ohl93]. GWAI-92 [Ohl93]. Gwangju [Ho12a, Ho12b, PZYS12].

H [HS05b, BH77]. Hagen [KIRK05, NS90]. Hagenberg [Alp11, AHK07, HRRY08, HNP12, KKMW07, Vor08, WL03, Win04]. Hague [GFV09]. HAID [AJB09, CKOB11, MSB12, MB06, NSFB10, OB07, PB08b]. Haifa [AVP07, BHKR11, Bih97, BZU07, CH09, DGR92, ELS12, FKK09, Gel08, Gru97, NZZ11, SZ92, UBW06, Yor08]. Hainan [DD06, STLK05]. HAIS [CAP08, CWO^{+}09, CRS10, CKW11a, CKW11b, CSA^{+}12a, CSA^{+}12b, RCS10]. Hakodate [AT07, BE06, BDDS07, BHJY07, DvEF06, FS07a, IST07, JSS08, JDMB08, KHSM^{+}08, MPR07, NVSB^{+}07, PZ07a, WPM07]. Haldia [JMT11]. Halevi [Gol11b]. Halifax [BJ12a, Deh97, DSZ07, JP07a, KC09, XCd03]. Hall [MSS04]. Halmstad [BS01a, BG03b, KL88]. Hamamatsu [MF01]. Hamburg [BPR05, BHJ^{+}11, BvdHPP09, Bra73, Bra80, BHH74, BRS09, BGW06, DFN10, GMS06, GKN03, HP12, HK96, LM98a, ND04, PRV11, Stu90, Vog88, ZK06, vM94]. Hamilton [DMRS07, Lau93a, MNS06]. Hammamet [DYCD09, Mel07, Per10b]. Hampshire [LG95]. Hampton [CMT02]. Handbook [Kou11]. Handheld [Gel99, TG00]. handicapped [ZBW94]. Handling [BH06c, DKRT06, RDKT01, SF08a].

Handwriting [DJ08]. Hangzhou [CDLZ06, CPZ07, CJ08b, FWY05, JRPP07, NPS06, PCH^{+}06, PAD^{+}06, PCK^{+}12, RSW12b, THL11, WCGB05, YLLZ04, ZYRH06]. Hannover [MSKY10, NKPH08]. Hanoi [BHH08, BOH^{+}12, GCSR07, HCL05, HZ08, HW05, JM02, Ngu06, RK09]. Haplotype [IWC04]. Haptic [AJB09, BMS01, CKOB11, MSB12, MB06, NSFB10, OB07, PB08b]. Haptics [Fer08, IS12a, IS12b, KvETvdH10b, KvETvdH10a]. Harbin [AGL^{+}07, BGG^{+}12, GLW^{+}12, LLM^{+}08, SLL^{+}06, ZLS^{+}06]. hardness [Sud95]. Hardware [CG09, CKLM09, GT03, GM06b, HHJ^{+}86, HSH08, JQ04, KKP02, KLZ07b, LK99, MS10a, MMC05, OR08, PV07, PT11, PS12c, RS05, Sha04, TTM10, WKP03, BD93, BP05, CE95, CB06, CYR02, HML97, KP99, KP01b, KNP01, KS97, Liu01, MM01, Mil00, MP93, MT99c, ST96, SMPU98, THT03, BL94, BHKR11, BZU07, CH09, ELS12, LB90, NZZ11, UBW06, Yor08]. Haring [HHG11]. Harrachov [vLIvdH^{+}07]. HART [Mal97a]. Hartmut [HL05, KMO^{+}05]. hashing [PS93]. Hasselt [CLS07a]. Hatfield [AS03, SC11]. Haus [ENR83]. Hausdorff [Ruc96]. Havana [NV01b, RSK08, SRS03, SL05]. Havel [VDD^{+}11]. Haven [Ban93]. Hawaii [DDI09, FR96, Fos99, JDMB10]. Hawthorne [ABRS06]. HBU [SGSV10, SL11, SdSMO12]. HCD [Kur09, Kur11]. HCI [Ayk07a, Ayk07b, Ayk09, Dai07, Duf07, Duf09, Duf11, FBSMS05, Har07, Har09, Har11a, HSL06, Jac07c, Jac07d, Jac07b, Jac07a, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, Kar09, Kur09, Kur11, LSHB07, Mar11a, Mar11b, OZ09, OZ11, Rau11, Rob11, SS09a, SS11c, SR07, SEG09, SF11, Sch07, SLH05, Shu07, Shu09, Shu11a, Shu11b, SS07c, SS07b, SS09c, SS11d, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, Chi03, CDM04, Hol07, Hol08, HM09, LHH10, MT^{+}03, Pat02, SLH04, TC95]. HCI-SWWA [MT^{+}03]. HCII [Ayk07a, Ayk07b]. HCITOCH [CFVCF^{+}12, FdCLN^{+}11]. HCIV [EDGP11]. HCSE [BFGL10, FP08, GHP^{+}08, WFB12]. Head [JPR04]. Health [AGB^{+}11, BR12a, Dai07, DPNM12, GMUW12, HMW^{+}08, Hol07, HS11, Kar09, LBM^{+}10, LW12, MKB^{+}09, OYM07, Ria08, Ria09, RtTMP10, RtTMP11, RtTM12, Rob11, TQYX11, Ste05, HLKX12]. Health-Care [RtTMP10, RtTMP11, RtTM12]. Heart [ADS09, CPR^{+}10, CKP^{+}12, FRSH05, KMC^{+}01, MA11, MKB^{+}09, SS07a, Mag03]. Hefei [DHX09, HZH05, WLC^{+}06]. Heidelberg [ASLM09, BB95, BZ83, Bor91, BK04b, BKO03, CHPW05, Ehr98, FMW01, GL10b, Gue01, HSJ07, Her92b, JO99, SSN08, Ste95, Ste94, Wet92]. Heights [Koz82, MMG^{+}01]. Heijen [KPS09]. Heinz [MMR93]. Held [AL11a, AI06, Ama08a, Ayk07a, Ayk07b, Ayk09, BH06b, Bar11, BBK12, Bir12, BHV11, CPR^{+}10, CKP^{+}12, CDCW09, CGMP11, Cas09a, CW09, Dai07, De 07b, DG12, Dro08, Duf07, Duf09, Duf11, DFGN12, DL07, EM10, FI08, FK12a, GO11, Gor10, GH07b, Gup10, Har07, Har09, Har11a, HK12, Hen08, HP06, HC11, Hof11, JCC^{+}12, Kar09, Kno11, KP09, KO07b, Kur09, Kur11, LMCH12, LSIT11, MDY^{+}10, MDHB11, Mar11a, Mar11b, MZ06b, O’B12, Ölv10, Ong10, OZ09, OZ11, PPCS12, RR08, Rau11, RKM^{+}12, Rob11, RT10, RKS^{+}12c, SS09a, SS11c, Sas05, Sch04b, SR07, SEG09, SF11, Sch07, Sei07, Sei12, Ses06, Shu07, Shu09, Shu11a, Shu11b, SS07c, SS07b, SS09c, SS11d, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f]. Held [Ste11d, Ste11c, SWSY11, TT10, VKG12, WYSS10, YLS^{+}12, YC11, YSL12, YHV12, dA09, dLZ12, dMS09, Ant97, AKL^{+}99, BCB99, BS76, Bod05, Boh75, BHR84, CRW97, Cer05, Cle99, Deg03, Dem80, Due04, Dur12, EN74, Est98, Fin99, GH03, Gor03, GS00b, HZ05, Han98, HWP01, Hed03, HAB98, HM01, Hor02, Hus01, Jäh99, JS97, JP04a, KG74, KS02c, Kos98, Kow01, LSW00, Le 02, Mai00, MY01, MM82, MSY00, NE02, NL99, Niv98, NvRSS11, PGM98, Pez03, Rol98, Rol99, Sag05, San01, Sel86, Smo00, Ste98, ŠPTM00, Swi99, SBH^{+}04, Tho99, Tiu00, Tob80, Wal04b, Wat00, WK02, WMS04, Wil01a]. Helenic [APS06]. Hellenic [DVVA08, KPK^{+}10, MPV12, VS02, VP04]. Helping [MKZ02, MKZB04, MKZK06, MKZK08, MKZK10a, MKZK10b, MKPZ12a, MKPZ12b]. Helsinki [EMT02a, EMT02b, FKS10, FK12b, HKT12, JN10, KS12a, KMC^{+}01, KOOL03, KC02, NU92, Oqu03, PRT06, YJKK06]. Henares [MRU12]. Her [ADM11]. Heraclion [LW95]. Heraklion [APS06, AGS^{+}11b, AGS^{+}11a, AZ11, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, BCP^{+}04, BDFS04, CCK95, CMV96, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, GCFA12, GPE05, GS06, Kut12a, Kut12b, LM06b, LMT^{+}04, MSS12, MS10b, MS10c, MS12b, NS98, NT03, SMBQ07, SCP^{+}12, TP09]. Heritage [CFVCF^{+}12, FdCLN^{+}11, IFGH10, IFL^{+}12, XCR11]. Hermann [CRS11]. Herrenalb [KK95, LK99]. Herrsching [MST95, SHL08]. Hersonissos [MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c]. HESSD [PVW10]. Heterogeneous [CMMR12, CJ06, Urs02, CJ05, PAC08]. HeteroPar [ADB^{+}12a, ADB^{+}12b, GVT^{+}11, LAF^{+}10]. HetsNets [BBK12]. Heuristic [Hel08]. Heuristics [SBH07, SBH09]. HI [APN08, BSvRW08, DSRW08, DKC10, HKM^{+}07, JSS08, LP08, RPR08, SPON08, WD10a, WD10b]. HiBB [ADB^{+}12a, ADB^{+}12b, GVT^{+}11]. hidden [dB93]. Hiding [BHJKPG06, BFSN10, CCJS07, FPCK11, FCDB07, KS09b, Shi06, Shi07, Shi09, Shi10, SEK^{+}11, Shi12, SK12, SSM08, And96, Auc98, Fri04, Mos01, Pet03, Pfi00]. hierarchical [Stu85, Beh03]. Hierarchies [MSS03b, BS76, She85]. High [CZ02, CDE^{+}07, DSS^{+}07, Kim05b, KN09b, Sch02, Ste07a, Ste09a, Ste11a, Ste11b, THX^{+}07, Yu02, BL92, BLW82, ELM05, Ham89, HR99a, M^{+}06, PPF^{+}00, Ste94, APBP07, BPSP06, BKS99, BP04b, BAWH00, CNmWH^{+}05, DDHP05, DPC^{+}07, DKS^{+}07, FLL03, GHR89, GH94a, GH94b, GK06, HS95b, HS97, HHW01, LJS08, LCHS96, ML04, Mei08, MCSN10, MPV01, Mue01, PDHS03, PAD^{+}08a, PDML11, PFD^{+}10, PCS^{+}07, PP03, Pol97, Pol99b, RPBP06, SPBP08, SKS02, SEO^{+}09, SBH98, Slo99, SDK^{+}08, VJKT00, VKV00, VJAA03, YRMD05, Yu02, ZCDT10, ZJS^{+}02]. High-Dimensional [Yu02, Yu02]. High-Level [Sch02, BL92, Mue01]. High-Performance [CDE^{+}07, DSS^{+}07, Ste07a, Ste09a, Ste11a, Ste11b, HR99a, LJS08, GH94a, GH94b, HS95b, HS97, HHW01, LCHS96, SBH98, Slo99]. high-speed [Ste94, FLL03]. Higher [BW03b, BNUW09, CMT02, Mah04, MMT08, NPW02, SB07, AH00, BDH^{+}99, BJ01b, GN98, GF97, HM05b, MSZ90, SBG04, vGH96, DHMM96, Hee94, JS94, MC94, SWAF95, Sha04]. Higher-Level [Mah04]. Higher-Order [NPW02, DHMM96, Hee94]. Highlands [HF97]. highways [BB96]. Hildesheim [Bes93]. Hill [CPC^{+}99, Der89, GF97]. Hilton [KJR05]. Hinterzarten [BRM06]. Hinxton [LNE06]. HIP [BL05a]. HIPC [BPSP06, APBP07, BP04b, RPBP06, SPBP08, MPV01, PP03, SKS02, VKV00]. HiPC’99 [BKS99]. HiPCoMB [THX^{+}07]. HiPEAC [CNmWH^{+}05, DKS^{+}07, PFD^{+}10, SEO^{+}09, SDK^{+}08]. HIPS [Mue01].

Hiroshima [AB04, GHI^{+}12, MPvH04]. Historic [JL11]. Historical [KD11]. History [GV08]. HM [AAB^{+}06, BBAB^{+}07, BBRS05, BBC^{+}08, BBD^{+}09, BBR^{+}10]. HOA [DHMM96, HHM97, Hee94]. Hobart [ShK06]. Hoboken [BB11, HMT10c]. Hoc [ASE^{+}07, BGW06, DPLW08, ENO12, Sch10a, WW07, ZOCJ07, BY06, CSJD06, CHPW05, JWH06, MTW05, SMCM07, SC05, KO07a, KR06, NBK04, NW10, PBK03, RGLA09, CSRS08, FLR11, LPR12]. Hoegaarden [SSv94]. Hofgeismar [KSW05]. HOL [NPW02]. HoloMAS [MPB05, MVC07, MSZ09, MVL11, MMV04, MŠKL02]. Holonic [MMV04, MPB05, MVC07, MSZ09, MVL11, MMV04, MPB05, MVC07, MSZ09, MVL11]. Homage [MFÁ^{+}09b]. Homburg [BH77, HS76]. Homes [AGB^{+}11, AN06, DPNM12, HMW^{+}08, LBM^{+}10, MKB^{+}09, OYM07]. Hong [BDDM12, BCF^{+}02, yCW96, CYGL04, CNX05, CSJD06, CT06, CTS10, CP06b, CWL01, CY12, Chi95, CP98b, CP98c, CBS^{+}11, CCAS12, DW07a, DY06b, FT04a, FKW08, FHW08, GPL99, HL00, cHPkC^{+}07, IMY^{+}07, IS98, IAL^{+}07, KWCW06a, KWCW06b, KWCW06c, KYT01, KZ10a, KFfKL11, LLCL05, LwCM00, LPC^{+}11, LMA09, LCB12, LYhL^{+}01, LCY03, LL06, LM08b, Luo11, Ng93, RWA06, SM11a, TLF01, TWYhL01, XYM^{+}07, XYZU11, YKF^{+}06, YKU11a, YKU11b, ZJ04a, ZJ06, Zha07, ZS10, ZYW08]. Hong-Kong [FHW08]. Honnef [BRR87a, BRR87b, CER79, DPP97, Gab97, HNW04, Len93, Nol81]. Honolulu [APN08, BSvRW08, DSRW08, DDI09, FR96, Fos99, HKM^{+}07, JSS08, JDMB10, LP08, RPR08, SPON08]. Honor [ACG^{+}07, ACÇ^{+}12, BCGY09, DHS10, HHG11, JLW07a, SPG10, HS05b, SKMR94, TJHS94]. Honour [ELLP12, KS02a, KS02b]. Hood [HSB^{+}10]. Hood/Portland [HSB^{+}10].

Horizons [LCY^{+}11, JLY03]. Hot [JMO10]. Houston [BBRM98, Cha05b, CMCS11, PV02, PCS^{+}07]. Houthalen [QV90]. HPC [M^{+}06, SMMG01]. HPC-GTP [M^{+}06]. HPCA [ZCDT10]. HPCC [GK06, GVT^{+}11, PCS^{+}07, YRMD05]. HPCF [GVT^{+}11]. HPCN [LCHS96, BAWH00, HHW01, Slo99]. HPCS [MCSN10]. HPCVirt [ADB^{+}12a, ADB^{+}12b]. HPF [PD96]. HPPC [ADB^{+}12a, ADB^{+}12b, BFT^{+}08, CAS^{+}09b, LAF^{+}10]. HPSS [ADB^{+}12a, ADB^{+}12b]. Hradec [FKM09]. HSCC [BBB07b, EM08, HT06a, MT09c, MT05b, AP04a, DSV01, LK00, MP03a, TG02]. HSCC’98 [HS98a]. HSCC’99 [VvS99]. HSI [SILS05, Chu03]. Hsin [CCL02]. Hsinchu [CL06a, CCH02]. HSNMC [FLL03, ML04]. HTN [Har11b, Ull08]. HTN-Planning-Based [Ull08]. Hualien [BCC^{+}10]. Huang [CWC^{+}07, DLW^{+}07]. Huangshan [DHP09, HJBF12, HMJG12]. Huangshi [CLKL09]. Huatulco [COMTLB12]. HUC [TG00]. HUC’99 [Gel99]. Hue [NLŚ10a, NLŚ10b]. Huerta [CFVV^{+}12, FdCLJ^{+}11]. HUG [JS94]. Huhehaote [QGZ03]. hulls [Knu92]. HUMACS [Ari02]. Human [BP07, BPR05, BD04, Cai08, CV04a, DGWB09, Duf07, Duf09, Duf11, EDGP11, EKF12, ERK07, EBAH08, FST01, FP08, GCK06, Hol07, Hol08, HM09, HS11, HSL06, HNPP07, IdKM06, KSFN08, Kim01b, KW10, LHH10, LN01b, MT09b, PB08a, SS09a, SS11c, SLH04, SS07c, SS07b, SS09c, SS11d, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, BZ83, Bra99, BMS01, BBB98, Fre03, GT90, GT05, LSW81, MJR04, Pad99, SLH05, TC95, WF98, WS02, BL05a, BPAB07a, BPAB07b, BUG93, BFGL10, BGU94, BGU95, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, Chi03, CFVCF^{+}12, CP05, EDGP11, ERK07, FdCLN^{+}11, GT93, GGK^{+}09a, GGK^{+}09b, Jac07c, Jac07d, Jac07b, Jac07a, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, KEM07]. Human [Kur09, Kur11, LK09a, LCHS08, LSHB07, PVW10, Pat02, SGSV10, SL11, SdSMO12, SS09a, SS11c, SS07c, SS07b, SS09c, SS11d, VU09, Vet11, WFB12]. Human-Centered [FP08, KSFN08, KEM07, WFB12]. Human-Centred [BFGL10]. Human-Computer [BD04, CV04a, DGWB09, EDGP11, EKF12, GCK06, Hol07, Hol08, HM09, HS11, HSL06, KW10, LHH10, LN01b, PB08a, SLH04, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, BMS01, BBB98, GT90, SLH05, TC95, WF98, WS02, BPAB07a, BPAB07b, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, Chi03, CFVCF^{+}12, FdCLN^{+}11, GGK^{+}09a, GGK^{+}09b, Jac07c, Jac07d, Jac07b, Jac07a, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, BUG93, BGU94, BGU95, CP05]. Human-Human [EBAH08]. Human-Machine [EBAH08]. Human-Oriented [BP07]. Human.Society [Chu03, SILS05, Kim01b]. Humans [SG07, SBF08, WK08]. Humlebæk [HK04]. Hunan [TNYL11]. Hungary [Ama08a, AFR^{+}04, BBvRL10, BDG04, BDG^{+}12b, BP03, BBT10, CGL^{+}09, CS10, CM09b, CÓ11, CTE^{+}06, CK93, CDG89, CVÉ07, DKP05, DT89, DZ03, DP10a, DH06, DKP00, Dro08, Esi93, EF03, Ési06, EVV^{+}11, FI08, FGTZ08, Fro02, FG95, GGS11, GR06a, GW12, Hen08, HY03, Hor06, HT06b, HZB07, HPZ10, JSO04, JFBH08, KKS11a, KKS11b, Knu01, KN83, KFM07, KVMH09, KKD04, KSTP99, KNP06, LRS03, Leu08, MML97, MRPM10, MTW05, MVA01, MWJ97, OV09, PAV^{+}10, PT08a, PJM^{+}08, PVP05, RR08, RAM05, RBC^{+}06, Rue93, SVV12, SV12, TT10, Vog01, Wei97, ZHP12].

Huntsville [WPD04]. Hurtado [MRU12]. HVC [BHKR11, BZU07, CH09, ELS12, NZZ11, Yor08]. HWTS [RKM^{+}12]. Hybrid [Ant95, Ant97, AKL^{+}99, Har11b, JLW07a, LHŚ11, LHKŚ12, SHŚ^{+}07, ZZ04b, Neb90, Rus05, AAB^{+}06, AHS96, AP04a, Ant97, AKL^{+}99, BBAB^{+}07, BBB07b, BBRS05, BBC^{+}08, BBD^{+}09, BBR^{+}10, CFK^{+}12, CAP08, CWO^{+}09, CRS10, CKW11a, CKW11b, CSA^{+}12a, CSA^{+}12b, DSV01, EM08, FKW08, Gro93, HS98a, HT06a, KFfKL11, LK00, MT09c, Mal97a, MP03a, MT05b, RCS10, TG02, TCLH10, VvS99, WFZL09, WS00]. Hyderabad [CV96, CChL^{+}10, DG04, GWK09, GG10, HNPP07, KH09, MS02b, MPV01, NNS06a, NNS06b, PR07, PP03, RS06, SP08, SAUL11, ZYRP10a, ZYRP10b]. Hyperedge [Hab92]. Hypermedia [BSS00, NKPH08, RA00, WAS06, BKS96, DN04, dBBC02, RTB02]. hypertext [NM97].

i.e [B^{+}12a, BI73, Bra73, GK90, HM04, Müh75, NP02, SF74, Sie75, Wan05b, Wil99]. I/IVR [CPSA10]. I/O [Bos96]. IA [HMW^{+}08, Len97]. IADS [CPGL05]. IAPR [BRP09, BV05b, CNA08, CSTD08, DRDKE11, DTMH10, EV07, GHI^{+}12, HWW^{+}10, JFT11, KSG^{+}09, dVLKR^{+}08, LWRdR11, MST12, PMS08, RSV07, SM06, SE10, ST12b, SKSA12, TEB09, YKF^{+}06, Ami98, CAD^{+}02, FIAP01, FCD^{+}04, HV03, IS98, wLN99]. IAPR-TC-15 [EV07, JFT11, TEB09]. Iasi [CP99, Gel10]. IAT [CvRVSN12]. IAT10 [GDG11]. IATA [AG98]. IATA’99 [Alb99]. IBERAMIA [GPAD08, KMS10, PDMFF12, SCR06, GRT02, LRG04, MS00c]. IBERAMIA-SBIA [SCR06, MS00c]. Iberian [AMPT09, MdlBP05a, MdlBP05b, MBMS07a, MBMS07b, PCPS03, VSH11]. Ibero [Coe98, GPAD08, KMS10, LRG04, MS00c, PDMFF12, SCR06, GRT02]. Iberoamerican [AMGJ12, BCE09, BC10b, MTOK06, RSK08, SK11, SRS03, STC04, SL05]. Iberoamericann [RMK07]. IBIOADT [IMW04]. IBM [BH77, BZ83, Hac75, HS76, HR91, MB82]. IBPRIA [MdlBP05a, MdlBP05b, AMPT09, MBMS07a, MBMS07b, VSH11, PCPS03]. ICA [AJRB09, DJAP07, REPH06, TCYZ12, VZM^{+}10, PP04]. ICA300 [XCHZ11a]. ICA3PP [BZ08a, HGZ05, HYPY10a, HYPY10b, HC09, JRPP07, XCHZ11b, XSA^{+}12a, XSA^{+}12b]. ICAD [YAKMJ10]. ICADL [BMC08, CCM^{+}04, CC12, CKH10, FNPW05, GCSR07, SHRM06, XCR11, LFK^{+}02, SZC^{+}03]. ICAI’96 [MWJ97]. ICAIS [Bou11]. ICAISC [RTZŻ06, RTZZ08, RST^{+}10b, RST^{+}10a, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, RSTZ04]. ICALP [AGK^{+}10, ADG^{+}08a, ADG^{+}08b, AHS11a, AHS11b, AMSM^{+}09a, AMSM^{+}09b, ACJT07, BPSW06a, BPSW06b, CIM^{+}05, CMPW12a, CMPW12b, AS94, BLPW03, DKLS04, FG95, LKC93, MM96, OSvL01, WMR00, WTM^{+}02]. ICALP’97 [DGMS97]. ICALP’98 [LSW98]. ICALP’99 [vWN99]. ICANN [APPE09a, APPE09b, DDI10a, DDI10b, DDI10c, DKOZ05b, DKOZ05a, HDGK11a, HDGK11b, KSDO06b, KSDO06a, KNK08a, KNK08b, VDÉ^{+}12b, VDÉ^{+}12a, dSADM07a, dSADM07b, DBH01, Dor02, GGHN97, Kay03, von96]. ICANN/ICONIP [Kay03]. ICANNGA [BDIR07a, BDIR07b, DLŠ11a, DLŠ11b, KTB09]. ICAPR [SSAP05a, SSAP05b, SMK01]. ICAPTN [CD05]. ICARIS [ATO^{+}09, BLJ08, BC06, CGK^{+}12, HMTH10, JPBT05, LNS11, dCVK07, NCBT04, TBH03]. ICAT [PCH^{+}06]. ICATPN [DT06a, KY07b, CK01, CR04b, DS98, EL02b, NS00, vdAB03]. ICATPN’97 [AB97]. ICATPN’99 [KD99]. ICB [LL07, TN09, ZJ06]. ICBA [ZJ04a]. ICCAL [Mau89, NS90, Tom92]. ICCBR [AW12, AW01, ABB99, AB03, BM10, LP97, MW09, MAR05b, RW11, WR07, VA95]. ICCBR-95 [VA95]. ICCBR-97 [LP97]. ICCBR-99 [ABB99]. ICCBSS [FP05, DG02, EW03, KP04]. ICCCI [JNH11a, JNH11b, NKC09, NHJ12a, NHJ12b, PCN10a, PCN10b, PCN10c]. ICCHP [MKZK06, MKZK08, MKZK10a, MKZK10b, MKPZ12a, MKPZ12b, MKZ02, MKZB04, ZBW94]. ICCI [AFK90, DFK91]. ICCL [BHJ^{+}11, HSSV12]. ICCL’98 [BCB99]. ICCMSN [PS09c]. ICCNMC [LZ05]. ICCPOL [LMA09, MSWZ06]. ICCS [ADJ^{+}01a, ADJ^{+}01b, AvAS06, AvASD06a, AvASD06b, AvASD06c, ANS^{+}09a, ANS^{+}09b, APHA11, BvADS08c, BvADS08a, BvADS08b, CFL10, EH08, PPH07, RDK09, SHØ06, SvADS07a, SvADS07b, SvADS07c, SvADS07d, SDM05, SvASD05a, SvASD05b, SvASD05c, BvASD04a, BvASD04b, BvASD04c, BvASD04d, DS01a, EEM96, Ell95, GM00, Luk97, PCA02, STDH02c, STDH02a, STDH02b, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, WPD04, dMLG03]. ICCS’93 [MMS93]. ICCS’94 [TDS94]. ICCS’98 [MC98]. ICCS’99 [CT99]. ICCSA [GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, GG07a, GG07b, GG07c, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, KKGT03c, KKGT03a, KKGT03b, LGK^{+}04c]. ICCV [SLH05, VHM07]. ICCVG [BKW09b, BTCW10a, BTCW10b, BTCW12]. ICDCIT [Cha05a, GM04a, JM07, JM10, MCK^{+}06, NO11, PA09, RR12]. ICDCN [AYV^{+}11, BDDM12, CDPT06, GWK09, KPSW10, RCJ^{+}08]. ICDEM [KA12]. ICDHM [Duf07, Duf09, Duf11]. ICDKE [X^{+}12a]. ICDM [Per05, Per06, Per07a, Per08, Per09a, Per10a, Per11a, Per12a]. ICDT [EL04, SS06a, AK90, AK97, Aus86, BH92, CLM03, GPV88, vdBV01]. ICDT’95 [GV95]. ICDT’99 [BB99]. ICE [MDH11a, Smi06]. ICEB [KZ10a]. ICEC [AFG^{+}11, HRC06, HMM12, KKKN05, MRN07, ND09, SS09d, YMHH10, Rau04b]. Iceland [ADG^{+}08a, ADG^{+}08b, BKR09, BFGL10, BD11, DR11, EH12, Fek08, FR11, Jón12a, Jón12b, KL96, LRH10, RBB^{+}06, VKMT11]. ICES [HSH08, KLZ07b, MMC05, TTM10, Liu01, Mil00, SMPU98, THT03]. ICES96 [HML97]. ICESS [LKK^{+}07, WCGB05, YZZ^{+}05]. IceTAL [LRH10]. ICFCA [DIP12, FR09, GG05b, KS07c, KS10c, MO08, MS06b, VJ11, Ekl04]. ICFEM [AT12, BC09b, BHLP07, DZ10b, LB05, LH06, LMA08, QQ11, DSB04, GM03]. ICGI [CCM08, PS04a, SKS^{+}06, SG10, AFvZ02, CO94, HS98b, MD96, Oli00]. ICGI-94 [CO94]. ICGI-96 [MD96]. ICGI-98 [HS98b]. ICGT [CEM^{+}06, Ehr04, EHRT08, ERRS10, EEKR12, CEKR02]. ICHIT [LHŚ11, LHKŚ12, SHŚ^{+}07]. ICHL [CFK^{+}12, FKW08, KFfKL11, TCLH10, WFZL09]. ICIAP [FSV09, MF11a, MF11b, RV05, BDV95, Del97a, Del97b]. ICIAR [CK04a, CK04b, CK06b, CK06a, CK08, CK10a, CK10b, CK12b, CK12c, KC05, KC07, KC09, KC11a, KC11b]. ICIC [HZH05, HLI06a, HLI06b, HLI06c, HHL07b, HHL07a, HWLJ08b, HWLJ08a, HJL^{+}09a, HJL^{+}09b, HZBF10, HZGZ10, HGBF12, HGGG12, HGPH12, HJBF12, HMJG12]. ICICA [LC11, LMC12, ZZLL10]. ICICS [CRW08, CY12, LQO04, NQL06, QMLW05, QIW07, QMW09, QSWL11, SQL10, DQBZ02, HOQ97, QOZ01, QGZ03]. ICICS’99 [VM99]. ICIDS [ALL^{+}10, IZP09, STA^{+}11, SS08b]. ICIRA [JLS11a, JLS11b, LDXZ10a, LDXZ10b, SRL12a, SRL12b, SRL12c, XXX^{+}09, XHXL08, XLHX08]. ICISC [Kim12, LC09, LH10, NR07, PC05, RL06, RN11, WK06, Kim02, LL03, LL04, Won01]. ICISC’99 [Son00]. ICISP [ELNM08, ELN^{+}10, EML^{+}12]. ICISS [BA06, JM05, JM11a, JM11b, MG07a, PG09, SP08, VG12]. ICITS [Des09, Feh11, Kur10, SN08, Smi12]. ICLA [BS11c, RS09a]. ICLP [DN07, DL04a, DPP95, ET06, GG05a, HW09, dlBP08, BG91b, Cod01, Pal03, Stu02]. ICM [BW03a]. ICMAI [AFS02]. ICMAS’96 [Wei97]. ICMB [Zha07, ZS10]. ICMCTA [Bar08a]. ICMI [HNPP07, TSG00]. ICML [ABF^{+}07]. ICMLC [YLWY06]. ICMS [FvdHJT10, IT06]. ICMT [CV11, HdL12, Pai09, TG10, VGP08]. ICN [LD05b, LD05a, Lor01a, Lor01b]. ICNC [JWbG^{+}06, JWG^{+}06, WCO05a, WCO05b, WCO05c]. ICOIN [CK06c, Kim05a, VFC08, Cho02a, Cho02b, Kah03, KG04]. ICONIP [HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, IDMY08a, IDMY08b, KWCW06a, KWCW06b, KWCW06c, KKC09a, KKC09b, LLC09a, LLC09b, LZK11a, LZK11b, LZK11c, WMB10b, WMB10a, Kay03, PKM04]. ICOODB [DZ10a, NG10]. ICOST [AGB^{+}11, DPNM12, HMW^{+}08, LBM^{+}10, MKB^{+}09, OYM07]. ICPIA [Nak92]. ICPO [BC95]. ICPR [ÜÇA10]. ICQT [CMS11, RST09, SPSV^{+}04, SSKR02, SRT06, SCKR03]. ICSC [Chi95, HL00]. ICSE [Con97, IJ06, SLL^{+}06, Som96, Est95, Lam96, TC95, WvdH03]. ICSEE [LFJI10a, LFJI10b, LJS^{+}10]. ICSI [TST10a, TST10b, TSCW11a, TSCW11b, TSJ12a, TSJ12b]. ICSNW [GHI^{+}06, BGKS04]. ICSOC [BCS09, BCT06, BKM08a, DL06b, DGT10c, DR09, FL09, GRB^{+}07, KMMN11, KLN07, LLTY12, MWYF10, MRY^{+}11, PJC^{+}12, OWPY03]. ICSOC-ServiceWave [BCS09]. ICSOC/ServiceWave [DGT10c]. ICSP [MDH11a, MYS10, WPR07, WPR08, WGMP09]. ICSR [EK09, GLCT10, GKC^{+}12, Mei08, Mor06, MBH^{+}11, Sch11, BK04a, Fra00, Gac02]. ICSR-6 [Fra00]. ICSR-7 [Gac02]. ICST [MSS11]. ICT [AKTT12, KT11, AKTT12, dSDL04, ST02]. ICT-GLOW [AKTT12, KT11]. ICTAC [BCC06, CDGW10, CP11, FHY08, HW05, JLW07b, LM09, LA05, RD12b]. ICTCS [CLP05, BL03, RRR01]. ICTIR [AC11, AKR^{+}09]. ICTL [GO94]. ICTSS [PSM10, WZ11]. ICUCT [SKCK07]. ICVGIP [KP06]. ICVR [Shu07]. ICVS [CD07, CDT11, FSP09, GVT08, Sub05, BST01, BST03, CH99, Cro03, SS01]. ICVW [CV08]. ICWE [ADP11, BFH07, BCKR10, BTT12, DF10, GGD09, HK12, LG05, Cue03, KFW04]. ICWL [LLCL05, LLLL08, LPC^{+}11, LZS^{+}08, LSL04, LLL06, LSW^{+}10, LCY^{+}11, PLK^{+}12, SLKL09, FCLL02, ZNCF03]. ICWS [JZ03]. ICWS-Europe [JZ03]. ID [AV07, FOGE^{+}09, RS09d, VDD^{+}11]. IDA [ARSB09, BSTL07, CAB10, FKP^{+}05, GBH11, HKT12, BLB^{+}03, HKB99, Hof01, LCB97]. IDA-97 [LCB97]. IDA-99 [HKB99]. IDCS [X^{+}12b]. IDEAL [CYBF06, CY09, FKLY08, FTC^{+}10, GHM05, LwCM00, YTC^{+}07, YWRS11, YCB12, LCY03, YEY04, YAF^{+}02]. Identification [Yal10, Bit88]. Identifying [DLN10]. Identity [AV07, CLS11, KL12, RS09d, SJDT08, JMK09]. IDGD [Ayk09, Rau11]. IDMS [BMO02, BMP96, DOS99, SvS00, She01, SW97]. IDMS/PROMS [BMO02]. IDMS’98 [GP98]. IDP [WZH^{+}12]. IEA [AE05, AD06, CHCA09, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, JDAW12, MMO^{+}11a, MMO^{+}11b, NBGA08, OA07, BR92, CHA03b, HA02, Ima99, LPA00, MdPA98, MVA01, OYA04, dPMA98]. IEA/AIE [AE05, AD06, CHCA09, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, JDAW12, MMO^{+}11a, MMO^{+}11b, NBGA08, OA07, BR92, CHA03b, HA02, Ima99, LPA00, MVA01, OYA04, MdPA98, dPMA98]. IEA/AIE-92 [BR92]. IEA/AIE-98 [MdPA98, dPMA98]. IEA/AIE-99 [Ima99]. IEC [TDB^{+}06, TDBP01]. IEEE [APS08, ASP01, ADM^{+}08, ASAM07, BG09, BFvdM10, BB00b, CGS07, DKK08, FLL03, GAA06, HJMP06, KMR^{+}05, KMF06, KSHM09, KPR07, LSHB07, LABM^{+}12, MMD05, MNPW07, MZC06, NCJK08, NSL09, PMÓ06, PAD08b, PB09, PISW05, RMDS10, SS05b, SvdMOP06, SGD09, VH04, YCC^{+}08, ZYW08, vdMBD08, AH02, ACK00, BK04b, CZ02, CCH02, FKB02, Fur95, FU96, LG94, LBP00, ML04, MA03a, SW04, SLC01, SS99, WGL96]. IEEE/ACM [BB00b]. IEEE/CS [LBP00]. IEEE/Nagoya [Fur95, FU96]. IEEE/Nagoya-University [Fur95]. IEKB [WZH^{+}12]. IFIP [AK10, ASE^{+}07, ASP01, Alo05, AS11b, AZ11, APP12, Atl08, BC09a, BBdB12, BPAB07a, BPAB07b, BA07, BdB08, BG09, BBK12, BKL^{+}12a, BKL^{+}12b, BPS06, BJ07, BP05, BAP^{+}05, BGR08, BD11, CKD11, CK12a, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CLWC08, CGMP11, CC07, CGS07, CCSW12, CP05, CFRR10, COB06, CBCGA12, DL06a, DPLW08, DSB10, DLNU11, DC12, DKK08, DV07a, DSZ10, DKU05, DFPS06, DPMP^{+}11a, DPMP^{+}11b, EDGP11, EM06, EK10, ES03, FR11, FJ10, FSTS09, GPAS^{+}06, GR12, GLIC10, GW06a, GH12b, GS08, GGK^{+}09a, GGK^{+}09b, GV09, GM10, HZ10, HJMP06, HHG11, HDLM09, HKM^{+}11, IR07, IS08, Jac04, JW05, JSWhT11, JRJ05, KD05, KK11e, KPR07, KH12, KA05, LLPH09, LN09, LM06b, LJJG07, Li11b, MBH^{+}09, MBVTY11, MT08c, MNW08, MW08, MPPU11]. IFIP [MCSY07, NPPDG06, NBM09, ONPR07, OS09, OSC^{+}08, PVW10, PZYS12, PAD08b, PVTG07, PSM10, PB09, PvS07, Pro11, QBY^{+}12, SNČ^{+}12a, STP^{+}10, SMBQ07, SJW^{+}12, SS05b, SO09, Smi05, SH09, SvdMOP06, SS06d, SZ05, SHYEF08, SHUH08, SV12, SSZ12, Tah09, TMG10, TMdB11, TMS12, TQYX11, TNP^{+}07, UDF06, VH04, Wan05a, Wim04, WCJS10, WZ11, vSH06, AAW04, AH02, ACK00, BCL03, Bis02, BGKS04, BK04b, CE95, CZ02, Cea76a, Cea76b, CGGO03, DO02, FKB02, GW96, GT03, GCC^{+}02b, GH04, Gue01, HW03, JGXC04, KHW03, LM98a, LK99, LPH00, LM03, LN01b, MM01, MNW01, MA03a, MP93, MKR^{+}04, MPS93, NNS03, NdG04, Omi00, PV02, PPF^{+}00, RS00, SW04, SS99, SDH03, STTP02, SC00, WSS04, Wim03, dFEN04, vLWH^{+}00]. IFIP-TC [BPS06, FSTS09]. IFIP-TC-11 [LM03]. IFIP-TC6 [ASE^{+}07, BAP^{+}05, DPLW08, CGGO03, GCC^{+}02b, MNW01, MKR^{+}04, Omi00, PPF^{+}00, STTP02, WSS04]. IFIP-TC6/European [Omi00, PPF^{+}00]. IFIP/ACM [Bis02, Gue01, SC00]. IFIP/GI [LM98a, LPH00]. IFIP/IEEE [ASP01, BG09, CGS07, DKK08, HJMP06, KPR07, PAD08b, PB09, SS05b, SvdMOP06, VH04, AH02, ACK00, BK04b, CZ02, FKB02, MA03a, SW04, SS99]. IFL [BGH06b, CHZ08, GH12a, GHMT05, HM11b, HZB07, MS11a, SC11, TMP04, AM02, HDC99, MK01, PA03a]. IFL’97 [CHD98]. IFL’99 [KC00]. IFM [BDS04, DG07a, DGLT12, LW09, MM10, RSvdP05, BPS02, GSS00]. Ifrance [MV12b]. IFSA [MCA^{+}07, BdBK03]. IGIS [Nie94]. IGKG [JPXS04b]. IH [BHJKPG06, BFSN10, CCJS07, FPCK11, Fri04, FCDB07, KS09b, SSM08, Mos01, Pet03]. IH’98 [Auc98]. IH’99 [Pfi00]. II [BGF06, BGHS05, CJ06, DvEF06, Dig11, JvdA09, Jon01, KKW12, LVS08, LOTY05, MBW08, Mül07b, NJA^{+}11, NVSB^{+}07, PCG^{+}10, PZH96, SCL^{+}06, WPM06, YK82, dSOSR06, AvLR91b, ADG^{+}08b, AHS11b, ANS04b, AMSM^{+}09b, AKM10b, ADB^{+}12b, ADJ^{+}01b, AvASD06a, APPE09b, ANS^{+}09b, AGS^{+}11a, Ant95, AHJ07b, AAH^{+}10a, AWA^{+}11b, AK99, Auc98, AOM07b, ADGM12b, Ayk07a, BBGS10b, BPAB07b, BHH04c, BS11e, BBP^{+}06b, BBP^{+}07b, BBP^{+}08b, BBP^{+}09a, BBP^{+}10b, BBP^{+}11b, BBP^{+}12b, BDIR07b, BF91, BKL^{+}12b, BTBP^{+}10b, BTCW10b, BRR87a, BHQ10b, BvASD04b, BvADS08b, BPSW06b, BGMST09b, BC98, BN98b, BC96c, CRJ11b, CK04b, CK06a, CK10b, CK12c, CGJ^{+}11b, CPFD^{+}03b, CZF10, Cea76b, CCD^{+}06b, CCP10b, Cho02b, CFG96, CCD^{+}12b, Coh96, CR73b, CRS10, CKW11b, CSA^{+}12b, CMPW12b, DGT10b, DGSM11b, DGM08b, DMP10b, DPB^{+}04b, DMLW11b, DBD^{+}10a, DBD^{+}11, DDI10b]. II [DO89b, DLŠ11b, DK02b, DPMP^{+}11b, EKLM87, Ekl94b, FSdlPT11b, FMP11b, F^{+}12a, FDC12b, FG04a, FTZ08b, FBHW00, FMC12b, GHJ06b, GPHF^{+}10b, GGK^{+}06c, GTW^{+}08, GTM10b, Gel11a, Gel12b, GGK^{+}05b, GG07b, GML^{+}08c, GTM^{+}09c, GL74b, GLC^{+}11b, GAY86, GGK^{+}09b, GHMV11b, HKW^{+}10, HLSZ11b, HSNJ02a, Ho12b, HDGK11b, HYPY10b, HLI06b, HCS11b, HZLL12b, IDMY08b, IS12b, Jac07d, Jac09d, Jac11b, JNR11b, JNHJ10b, JNH11b, JLS11b, JNPV10b, JWG^{+}06, Jón12b, KS03a, KS02b, KC11b, KKS11b, KKS11d, KWCW06b, KILW10b, KW98, KH11b, KSDO06a, KDH^{+}11b, KKC09b, KNK08b, KKA05, Kut12b, LGK^{+}04a, LNS06b, LTL^{+}11b, gLPZ^{+}12b, LBP06b, LLC09b, LSDN04b, LFJI10b, LSTZ12b, LFH^{+}07b, LDXZ10b, LZP^{+}11b, Lor01b, LD05a, LHJ08b, LZK11b, LGR^{+}04, MF11b, Mar11b, MS10c, MS12b, MŠKL02, MdlBP05b, MBMS07b, MT06b, MTH06b, MT07c, MTH07b]. II [MT08b, MDH09b, MDH10b, MDH^{+}11c, MMO^{+}11b, MAFS08b, MKZK10b, MKPZ12b, MP01b, MÁ05a, MÁ07b, MFÁ^{+}09a, MDPQA12b, MSY00, MGI^{+}11b, MGM^{+}12b, NNS06b, NHJ04b, NLŚ10b, NK10b, NKJ11b, NHJ12b, ORS89b, ORB^{+}09, OPT01, OF15, PM04b, PCME09, PCN10b, PCN12b, PSDS11b, PSPH^{+}10b, PSD^{+}05, PSM^{+}07, PRH10b, PT91b, PZ05, QLK^{+}11, RA06b, RDPMA^{+}11b, RST^{+}10a, RKS^{+}12b, SS09a, SS11c, SCKR10b, SJHJ10b, SvADS07b, Shi07, Shu11b, SBD03, SAG10b, SYP^{+}05, STDH02a, SAB^{+}03b, SS07b, SBJ^{+}05, Ste09a, Ste07b, Ste09d, Ste11f, SRL12b, SZT^{+}08b, SvASD05b, TST10b, TSCW11b, TCHB12a, TSJ12b, TvdBI02, TKCW11b, TGM^{+}10b, VRHJ09b, VDÉ^{+}12a, WLY05b, WYZ^{+}06b, WDGL10b, WYP12b, WMB10a, WMT96, WWC02, WD10b, WDKW10b, WDKW12b, XCHZ11a, XSA^{+}12b, XLHX08, YKKZ07b, YHR^{+}09b, YWG04b, YHZ09b, YKU11b, ZYRP10b, ZRP^{+}11b, ZiTM10b, ZLK10b, dBNT87b, dBM10b]. II [dLGR04, dSADM07b, dPMA98]. II/JURIX [CPSA10]. IICS [BRU06a, BU01, BBB06, UBM02, BHU03]. IIDB [GHI^{+}06]. IIHA [GHI^{+}06]. III [AGM05, BEOT06, FLVS11, HENR03, HKW11a, KBHSW06, MMP03, PPCS12, SLL^{+}07, SPON08, TNGK08, VS05, WPM07, vdBK05, ANS04c, AvASD06b, AHS96, AHJ07c, ADGM12c, BBGS10c, BBP^{+}10c, BvASD04c, BvADS08c, Buh98, BE01, CGJ^{+}11c, CGLR05, DMP10c, DSM94, DMLW11c, DDI10c, DC01, FMP11c, FLP^{+}12c, FTZ08c, Fre03, FMC12c, GHJ06c, GPHF^{+}10c, GGK^{+}06b, GT09a, GGK^{+}05c, GG07c, GOW03, GHMV11c, HSNJ02b, HLI06a, HZLL12c, Jac07b, Jac09a, Jac11c, JBK09, JNPV10c, KKS11e, KWCW06c, KSW99, KDH^{+}11c, LGK^{+}04b, LBP06c, LJS^{+}10, LFH^{+}07c, LZP^{+}11c, LHJ08c, LZK11c, MWJ97, MGI^{+}11c, MGM^{+}12c, NHJ04c, Ngu11a, PM04c, PCMC09, PCN10c, PCN12c, PS05a, PTZ03, PMGO05, RA07a, SJHJ10c, SvADS07c, STDH02b, SAB^{+}03c, Ste11a, Ste07c, Ste09c, Ste11d, SRL12c, SvASD05c, TGM^{+}10c, VPK00, WLY05c, YHZ09c, ZiTM10c, ZS05a, dLGR05]. IIST [AM03]. IJCAI [HNPP07, LSJ06, MA05c, PAV^{+}10, WS96, DER94, FO95, Ral94, RM97a, WMT96]. IJCAI’97 [RS99a]. IJCAR [ABD08, FS06, GH10, GMS12, BR04, GLN01]. IJCCGGT [ABK05]. IJCNLP [DWSK05, STLK05]. ILC [ABGL09]. Illinois [Cow77, Lom93, LO88, Mar93]. Ilmenau [BU01]. ILP [BRST08, De 10a, FL11, KP05, MOTN07, MTNL12, ŽL08, CF00b, DF99, HY03, LD97, MS03b, Mug97, Pag98, RS01, CKS04]. ILP-96 [Mug97]. ILP-97 [LD97]. ILP-98 [Pag98]. ILP-99 [DF99]. ILPS [Fre98]. IMA [Che11, Gal07, HMS09, MBS05a, MSW07, Par09, Sma05, Boy95, Dar97, Hon01, Pat03, Wal99a, WM03]. IMACC [Che11]. Image [ABB^{+}11, AMGJ12, AMPT09, AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, BHH04b, BHH04c, BBA12, BCE09, Beh03, BS06b, BIK01, BC10b, BBH08, CL06a, DYCD09, DCFR12, DLN10, DK02a, DK02b, DG05a, DFGN12, EP03a, EP03b, FFL^{+}12, FMP11a, FMP11b, FMP11c, FDL10, GMV03, Ho12a, Ho12b, JNPV10a, JNPV10b, JNPV10c, KP06, KŽ04, LNS06a, LNS06b, LPZ12, LSIT11, MDHB11, MdlBP05a, MdlBP05b, MBMS07a, MBMS07b, MTOK06, MR07a, MAFS08a, MAFS08b, PCPS03, PLG06, REF^{+}06, RMK07, RSK08, SK11, SRS03, SPO11, SKK04, SRM^{+}11, VSH11, WHL09, WB09, WR09, YHR^{+}09a, YHR^{+}09b, YLS^{+}12, ZJL06, BK81, DDJ00, DG05b, Jah93, Len90, MB97, Nak92, NV01a, Per05, Per06, PVC^{+}99, Rei93, STC04, SL05, TC99, Tuc01, WC98, ZLJ05, ZVU99, BHL^{+}03, BG03b, BDV95, CK04a, CK04b, CK06b, CK06a]. Image [CK08, CK10a, CK10b, CK12b, CK12c, Chi95, Del97a, Del97b, ELNM08, ELN^{+}10, EML^{+}12, EKO^{+}04, EP07, FSV09, HK11, KPK05, KC05, KC07, KC09, KC11a, KC11b, LLC^{+}05, LSE02, MF11a, MF11b, RV05, SHJ09, SNSR06, KHG01]. Image-Guided [MDHB11]. Imagery [BRP09, CSTD08, DRDKE11, KNP06, NSS03, AF97, ADL05, BCP99, BNS00, BLV02, MMU96]. Images [BBH^{+}10, JP09a, KKH07, PS07a, PS08a, PCG^{+}05b, PW03, RDPMA^{+}11a, RDPMA^{+}11b, ÜÇA10, AS99, CK93, GP05, HS95c, KvG98, Pol01, Ska01, SL99b, SDP97]. Imaging [ADS09, CKP^{+}12, CS05b, DSL08, FRSH05, GML^{+}08a, KL07b, KMC^{+}01, KK09a, LEP^{+}10, MDY^{+}10, MDHB11, MBG12, MLTC11, MA11, PPM09, SS07a, STT11, SWSY11, SH11, TN03, TST09, WYSS10, YJ04, YJS^{+}06, YC11, YSL12, YHV12, AKR03, BG93a, CH91, DG97, HAB98, IL01, KSTP99, Mag03]. Immune [ATO^{+}09, AMNT06, BLJ08, BC06, CGK^{+}12, HMTH10, LNS11, NCBT04, TBH03, dCVK07, JPBT05]. Impact [CR05b, Rüt98, DPNM12]. imperfect [BLS02]. Imperial [Sha86]. Implementation [BAdVR08, CBF09, DZGS06, FD03, HP02a, HOC07, HHLR08, Tah00, Vaz99, dVGR04, BJ01a, BW92, BP93, Buc90, CL96, CMZ99, DLM89, DM90, Dor95, Ell95, Fit93, FLS05, HP94, KDS01, LSDP81, LPSD83, MW91, Mio90, Mio93, PN93, RWY97, She85, Tah01, WF77, WY98, AM02, BMCCM11, BGH06b, CM03a, CHZ08, CHD98, DOSY05, DS11c, FLS06, GH12a, GHMT05, HM11b, HDC99, HŽ07, HZB07, ID03, IY06, IR08, Klu97, KC00, Man09, MK01, MS11a, MR12c, PA03a, SC11, Sch01b, TMP04, WW03, YP01, BC89, EK77, Mar80]. Implementations [BGS^{+}03, GHK97, HS95a, KS96, Tha84, BCK08]. Implemented [Ull08]. Implementing [BJ01a, CMZ99, RWY97, SBH07, SBH09, WY98, FNPW05, Hal96]. implications [Oss99]. IMPROVE [NM08].

Improvement

[ABMM07, BVO10, BI04, BKS05, COBV11, DJJ12, JS08, MV06, MA07c, OKS02, RRM06, BO00b, BKS01, Din04, RAM05]. Improving [Bee02, Han02]. IMS [CKO01]. IMTCI [BMN05]. IMWS [KRMM^{+}02]. INAP [AS11a, SHGB05, SHW09, UWB^{+}06, Bar03]. INBAST [MDH11a]. Incentives [SHS10]. Incheon [KGV^{+}09, NJHJ08]. including [GHK97, ZVU99]. Inclusion [HM09]. Incomplete [AGT98, Sch98b, Gra91]. Inconsistency [BHS05]. increasing [OZ01]. incremental [Ram96, Amt99]. Independence [LBM^{+}10]. independent [Wal99b, AJRB09, DJAP07, PP04, REPH06]. Index [MZ98, Jür02]. Indexed [Gib12]. Indexing [Yu02, Gra96]. India [AR99, AS02a, AYV^{+}11, APBP07, AKG06, AR98, AP07, Asa06, BPSP06, BA06, BKS99, BS11c, BL06a, BC11a, Bha95, BN91b, BP04b, BL07b, BHM03, BB00b, CV04b, Cha05a, CM12, CV96, CDPT06, CMM^{+}09a, CChL^{+}10, CD00b, CRD08, CFRR10, DB02, DD03, DSDS04, DG04, DU09, DBC^{+}10, DLW12, DBM09, GN12a, GWK09, Gel12a, Gel12b, GM04a, GDP07, GG10, GM10, HMV01, HKP08, HNPP07, JM05, JM11a, JM07, JM10, Jha10, JM11b, JM03, JS84, Jos00, JMT11, KP06, KA12, KPSW10, KP00, KK11b, KH09, KMMP12, LM04, Mad89, MCK^{+}06, Mah85, MMV06, MG07a, MS02b, MPV01, NNS06a, NNS06b, NO11, Nor86, Nor87, NK88, Nor90, PS02a, PKM04, PBB05, PKKG06, Pal03, PR07, PDSD10, PSDS11a, PSDS11b, PDSN12, PA09, PP03, PG09, RCC12, Ram08, RRR99, RS10, RCA90]. India [RS06, RS09a, RR12, RS97a, RD01, RCJ^{+}08, RPBP06, RO00b, RM04b, Roy05, RS09c, RD12b, SPBP08, SKS02, SCH^{+}09a, Sar03, SP08, Shy92, Shy93, SAUL11, SRY07, SB12b, Thi94, Thi95, TPCB12, VKV00, VG12, YAKM^{+}12, ZYRP10a, ZYRP10b]. Indian [BS11c, KP06, RS09a]. Indiana [Ang97, AKL^{+}99]. Indianapolis [Ang97, Lei95]. Indias [CBGCCZ12, PDMFF12]. Indies [DR06, Fra01, Fra99, Hir97, Hir98, PH04, Syv02, TTF07, Wri03]. Indigenous [SZC^{+}03]. Individual [BSPB07]. Indo [KMMP12]. Indo-Japan [KMMP12]. INDOCRYPT [BL06a, BC11a, CV04b, CRD08, GN12a, GG10, MMV06, RS09c, SRY07, JM03, MS02b, RD01, RO00b]. Indonesia [ABK05, BMC08, FV10]. Indonesia-Japan [ABK05]. Inducing [AGT98]. Induction [SW10]. Inductive [AJ94, BB06, BRM06, DFKM08, DS07, MLK04, SKP10, BJS93, DJS91, GS05, Jan87, Jan89, Jan92, NCW97, BRST08, CKS04, CF00b, De 10a, DF99, FL11, HY03, KP05, LD97, MS03b, Mug97, MOTN07, MTNL12, Pag98, RS01, Sch03b, ŽL08]. Industrial [AE05, AD06, ACJ09, BF10, CHCA09, CHA03b, CF09b, CCSW12, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, HA02, Ima99, JDAW12, KS07a, KR10b, LM08a, LPA00, MMV04, MPB05, MVC07, MSZ09, MVL11, MMO^{+}11a, MMO^{+}11b, MVA01, NBGA08, OA07, OYA04, Per06, Per07a, Per08, Per09a, Per10a, Per11a, Per12a, PNB04, PR98a, SS11b, SNZ08, SP12, ABL96, Ber99, FJL97, GW96, Kåg98, NDB94, NSM00, Per05, PGV93, SVV12, Waś96, WL93, BR92, MdPA98, dPMA98]. industrial-strength [Waś96, WL93]. Industry [BCT02a, BMUV10, BP12a, ECB06, PS08a, SB12a, SCH04c, UMN^{+}08, SMMG01, Rei03]. iNetSec [CKD11, CK12a]. INEX [FLMS05, FLMK06, FLT07, FKLT08, GKT09, GKT10, GKST11, GKS12]. Inference [AFvZ02, AJ94, BPCS06, BMS04b, CCM08, CPR12, GJN10, IWC04, SKS^{+}06, SG10, SHL08, CO94, Cip95, FW92, HNM02, HS98b, Jan87, Jan89, Jan92, KIRK05, MD96, Oli00, PS04a, DF02]. Inferences [Küs01]. Infinite [GTW02, Pup10, Bur97, NP85]. infinite-state [Bur97]. infinity [MvOvRdV05, ABGL09]. Informatics [BKLP11, BKLR12, BMW10, BH05, BR07b, BR09, BZ03, CWZ^{+}11a, CWYC12, CMZL04, CWY^{+}06, CYCL09, CChL^{+}10, CVV12, CHK06, DSZ^{+}11, DL81, FC04, FB12, GML^{+}08a, GS01, GP02, HJW00, HKV10, HLCZ11, KM11, KMR^{+}05, KLP10, LBNF08, LW12, LB78, LO10, MZC06, Mit05, Mit06, MS08a, OALZ^{+}08, PR01, PVV10, Rov98, Sam76, Sim92, VV07, Win11, YZC^{+}07, YCC^{+}08, YSP^{+}10, ZRP^{+}09, ZRP^{+}11a, ZRP^{+}11b, ZTTY12, ZGK^{+}07, ZLY^{+}07a, ZLLC09, ABD04, BYGP95, BSW95b, BBP96, BBZ00, BBZ01, Che03a, GPV00, JKB96, LM98c, MEK03, PTB99, PJ97, Raj02, vLWH^{+}00, KM11, Mit06, MS08a, Wil01b]. Informatik [BH77, Hac75, HS76]. Information [AB10, Abe10, AKK04, APR^{+}06, Abr07, AFP^{+}10, ABC^{+}06, ANS04a, ANS04b, ANS04c, ACR07, AC11, ATM09, AFGvE10, AFGvE11, ABJ02, AGT98, AHJ07a, AHJ07b, AHJ07c, ALZ11, ACJ08, ADK04, AKR^{+}09, BYdVZ^{+}12, BW03a, BMS^{+}08a, BGG^{+}12, BdQ11, BH06c, BPR02, BL08, BCG^{+}07, BDG04, BRL08, Bha00, Bha05, Bha07, BB03, BS10, BMYZ91, BMYZ95, BBMSD09, BHL11, BIMW12, BKL^{+}12c, Boy01, BGHS05, BB02, BMC08, BkHJ^{+}04, CBGCCZ12, CCS04, CT06, CFM09, CIT11, CWC^{+}07, CL10, Che99a, CLZ^{+}09, Che10, CTYG10, CTS10, CP06b, CKLN10, CBS^{+}11, CW94b, CFG^{+}11, Col07, CCSW12, CMB^{+}08, CDM04, CR05b, CFS06, CFL10, CHR10, Dat09, DdQ10, DGJ^{+}05, DPT12, DTW03, DLW^{+}07, DP06, EKS10, EHKP05, EBT11, EM02, EGMW11, ES02, EBT07, EKM06, EH12, FRvKS10, FWY05, FGH98, FKK09]. Information [FWZH06, FX08, FG06, FGK^{+}11, FLMK06, FLT07, GHJ06a, GHJ06b, GHJ06c, GKL09, GLW^{+}12, GHSW04, GZ09, GLC^{+}11a, GLC^{+}11b, GJS08, GSS11, GMV09b, GKMN10, GHK^{+}10, HG02, HTW02, HY12, HRdV11, HKI08, HZK08, HLKX12, HC11, HN12, HK05b, HdQ08, HRM^{+}10, HMMW10, HCDL09, HNS^{+}12, HXC^{+}08, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, HKH10, IMW04, IMK06, INR07, IAL^{+}07, IDMY08a, IDMY08b, IN11, JvdA09, JMT11, KMTW03, KL07a, KSS08, KTH09, KLM^{+}07, KHJ05a, KHJ05b, KHJ05c, KSB10, KMSB11, KM96, hKA10, hKhLKŚ10, hKAS^{+}11, KWCW06a, KWCW06b, KWCW06c, KOS02, KOOL03, KOK04, KRP06, KHPS07, KPP08, KLLP12, KBHSW06, KHSM^{+}08, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KLPP03, KFMM06, KKC09a, KKC09b, KY11, KS04c, KOS07, KWT03, Kur07, KŽ10b, LO02, LC06, Lai03, LMR^{+}06, LOX99, Lau00, Lee04, LYMM05, hLhKcFŚ09, LSL^{+}09]. Information [LW11, LHŚ11, LHKŚ12, LdQ07, LLC09a, LLC09b, LSY^{+}06, LLP^{+}07, LWF04, LLM^{+}08, LXH^{+}12, LP10b, LSK08, LLWL09, LK07, LK09b, Lop06, LH08, LSW12, LHJ08a, LHJ08b, LHJ08c, LZK11a, LZK11b, LZK11c, LS12b, LCDT08, MOP^{+}08, MN02a, MPS06, Mat09, MF08, MT04b, MSW02, MM04, MKR07, MMM05, MBJ^{+}07, MHW12, MR11, Mue06, MWK^{+}09, MMM11, MZWZ05, NdMO03, Nau02, NHJ04a, NHJ04b, NHJ04c, NLKJ06, NKN^{+}05, NLŚ10a, NLŚ10b, NKJ11a, NKJ11b, OiDP98, Oka00, OdQ12, OKdQ09, PKM04, PHJ03b, PHJ03a, PST11, PCN12a, PCN12b, PCN12c, PCA^{+}09, PSE10, Paz03, PRK12, PR05, Per10b, PS04b, PBGK03, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, PMWO02, Pie08, PZ07b, QLXL07, QLK^{+}10, QLK^{+}11, QBY^{+}12, RFBW12, RS00, RMFG06, RG03, RB10, Roy05, SL12, SSO^{+}11, SS09a, SS11c, Sch98b, Sch74, SVS08]. Information [STT04, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, SG09, ST02, SPÖ^{+}10, SF08b, SS07c, SS07b, SS09c, SS11d, SLSW12, Som09, Sta99, SS04, SHŚ^{+}07, TM09, TBE08, TFK11, TK05, TFE^{+}09, TKSN11, TQYX11, Urs02, Vää99, VY93, VRHJ09a, VRHJ09b, VU09, VLY09, WGLL10, WLO^{+}11, WS12, WJL^{+}12, WT07, WHG07, Whi00, WDKK07, WMB10b, WMB10a, XW11, XKGS12, YBM87, Yak00, Yak02, Yak04, YN06, YŞ03, YGGÖ05, YSR^{+}06, YKL06, Zhe02, ZSP^{+}04, ZYRH06, ZBY07, vEGW09, vdBK05, vM94, ACP01, AFK90, ABS91, AM04, Ari02, Aso04, BJM^{+}05, BDGM05, Bat01, BCBG05, BB96, Bha02, BP95, BKM01, BL06c, CC05, CE01, CF80, CCL02, CCH02, CM05a, CN05, CMV96, CGvR02, DG04, DFK91, DGN01, DH06, DL81, ERW99, EM03, EFM04, ES00b, FC93, FK95a, FM99, FLMS05, Gra91]. information [HNR05, HF97, HK05c, HS99, ILR95, IS98, JOD98, JLY03, JO99, KK97, Kay03, KIRK05, KHJ05d, KW98, KSW99, KK00, KZ01, KBEP03, KBC05, Leu97, LV05, LMV98, LFL05, Lou92, LZ00, Mon01, NOOS05, Nie94, OP97, Ooi90, PZ92, PF05, Paz97, PW92b, PT98, Pet01, PBGK02, PGBK04, PT99, RBC93, SS83, ST00, SS91, Sch97b, Seb03, SLC01, Smi98, SSB90, SKB98, Str99, ŠPTM00, TPB79, Vck99, Vis04, WYL01, W^{+}12, WB00, Wan93, WBW94, ZHF04, van97b, AJNJ12, ALM95, And96, AZ11, APP12, Auc98, AJN12, BA06, BLW09, BYLJ11, BW11, BHJKPG06, BG93a, BS02a, BSN06, Bha95, BKLP11, BKLR12, BFSN10, BCD00, BM03c, BG05, BN09b, BTMI11, CCJS07, CFH^{+}12, CY03, CG02, CDLZ06, CRW08, CMS08a, CKLN10, CY12, Cho02a, Cho02b, CK06c, CFG96, CS05b, CSY09]. Information [CY11b, CBS02, CH91, CR05b, DMO98, DF01b, DB98a, DW07a, DQBZ02, DBPZ05, Des09, DG97, Feh11, FLY06, FPCK11, FNY93, Fri04, FCDB07, GLMP07, GF12a, GSP01, GS94, HOQ97, HS11, HNS^{+}12, IL01, JM05, JM11a, JM11b, JY12, Kah03, KG04, Kar91, KL07b, KLB^{+}06, KS09b, KSFN08, KLP10, Kim02, Kim05a, KYL07, Kim12, KT11, KSTP99, Kur10, KDWW10, LYL11, LZL11, Lau12, LL03, LYMM05, LYY07, LSL^{+}09, LC09, LH10, LY12, LLM^{+}08, LL04, LY05, LYL06, LC11, LMC12, LB78, LST^{+}12, LQO04, MZ99, MBH^{+}09, MG07a, Mos01, MSS08, MZWZ05, NR07, NJ10, NLKJ06, NQL06, OSC^{+}08, PH11a, PC05, Paz97, Paz99, Paz03, PYLW08, Pet03, Pfi00, PS96b, PSNS99, POS00, PGD07, PK12, PG09, PPM09, QOZ01, QGZ03, QMLW05, QIW07, QMW09, QSWL11, Ren02, RL06, RN11, RSW12b, SNS03]. Information [SN08, SSO^{+}11, SYMA09, STP^{+}10, SMBQ07, SP08, Smi12, SSM08, Som09, Son00, SKY06, SQL10, SH10, SMS12, SV12, SH11, TN03, TY11, VPM97, VM99, VM01, VFC08, VG12, WPV04, Won01, WK06, WLRL08, YY09, YLL09, ZZ04a, ZG03, ZLDB05, ZZLL10]. Informationszentren [Sch74]. InfoSecHiComNet [JMT11]. InfraSec [DFR02]. Infrastructure [AL06, BB07, BSW12, CCBW11, HS08b, KGL04, LSF07, LSW12, MMK08, SP09, SNČ^{+}12a, SG09, SD10, WR01, CZ05, KRMM^{+}02, KMN94, DFR02, WR01]. Infrastructures [ABD^{+}12, CL11, Lop06, LH08, MP10b, PNPP12, RB10, SKM07, XW11, KMTU03]. Initiative [CFH^{+}12, FLMK06, FLT07, FKLT08, GKT09, GKT10, GKS12, FLMS05, MO92a]. Inititative [GKST11]. INLG [BEP04]. Innovation [JNS05a, WKS^{+}10]. Innovations [WKB04, ZRP^{+}11a, ZRP^{+}11b, AE05, OYA04, PM08]. Innovative [DO89a, DO89b, EEK^{+}87, EKLM87, BRU06a, BU01, BHU03, BBB06, HRR^{+}06, NT06, THR03, UBM02]. INNS [MST12]. Innsbruck [BL05b, FKBA12, FKM07, KASS07, MPdSM07, Pfi01, PMDA96]. INOC [PRV11]. Input [Bre97]. Input/output [Bre97]. INRECA [Ber99]. INRIA [BV92]. Inscrypt [BYLJ11, LYL11, LYL06, PYLW08, YLL09]. Insights [SO99]. Inspiration [SHØ06].

Inspirations [DKOZ05a]. Inspired

[GTP12, IMW04, IMK06, LYCV08, MRC12, MSA99b, MP01b, MÁ07b, SKRZ04, WAW01, ZHLW12, CSPC09, HGPH12, LFIM07, LYCV08, MÁ07a]. Instantiation [Wil91]. Instinct [Cai11]. Institut [BHR84]. Institute [AM03, BB89, NS81]. Institutions [BPD06, CvRVSN12, DFPV11, HMBD09, NVSB^{+}07, PAV^{+}10, SPON08]. instruction [BHH74, DF96]. Instruments [dSOSR04, dSOSR06, BND01]. Integer [AG01, BC95, BN04, BBRM98, CS02b, CBW99, CMQ96, ES10, FW07, GW11b, JK05b, LPR08, Wal99b]. Integrated [Nau02, SPB00, ALM95, HNR05, Nag96, RD84, AS99, AGCP^{+}11, AS07c, BG09, BDS04, BPS02, CM03b, DG07a, DGLT12, GSS00, HAB02b, LW09, MPK04, MM10, MvL10, PVV05, RSvdP05, SM09, THL11, VAM06, vLS11]. Integrating [FST01, FBHW00, OO12, HR91, SKB98, Str99, CC95]. Integration [BCBF08, BR12b, CBT07, GS06, GB07, LK10, LNE06, LR05b, PMH09, PF07b, Rah04, RR04a, Bol99, Bre03, Dav81, Gue05, Kar91, KU94, PCA02, Vis04, AB11, BM05b, BS06a, BJP12, EDD04, LMT10, PT08b, VW07, vHH09, Ger04, KS95, LCS91]. Integrity [BP95, BP95]. INTELLCOMM [AAW04]. intellectual [GT05]. Intelligence [ACvLdR03, ACdR^{+}08, ABG09a, AL04, AE05, AD06, AFS02, AIL12, APS06, AP11, BV02, BE11, BM05a, BPP91, BB90, BP07, BS11d, BS11e, BAHH07, Ber08, BVAT12, BBKM11, BR02, BV12, BMN05, BCC07, Bra09b, BJ01c, BF04, BCDN07, BL11, CPS05, CSPC09, CRJ11a, CRJ11b, Cai05, CA06, CLKL09, CHKL10, CBC^{+}02, CIW06, CT03, CD00a, CVV08, CMM^{+}09a, CMZL04, CHCA09, CCK95, CV10, CSAH09, CUdlC04, CAP08, CWO^{+}09, CRS10, dRCVT11, CF10, CFL10, DVVA08, Del00, DWWL09, DMLW11a, DMLW11b, DMLW11c, DBB^{+}08a, DV07b, DD10, DBHS10, DPT08, DKB03, DGB^{+}06, DBB^{+}08b, DBD^{+}10b, DBB^{+}12b, EL07, ETG02, ED06b, FK10, FMC04, FV10, Fil91, FvdHKL06, FKS07, GJ09, GF03, GPAD08, GdATM05, GRG06, GM07a, GM08a, GDP07, GK12, GKN03, HWP01, HKI^{+}09, HA02, HV06, HC11, HBE07]. Intelligence [HZ08, HLW08, HYMZ11, HLI06a, HLI06b, HHL07a, HNPP07, HWLJ08a, HJL^{+}09b, HZGZ10, HGGG12, HKH10, IPSK04, Ima99, IMY^{+}07, IS02, JN10, JKL02, JNH11a, JNH11b, JDAW12, KLZ07a, KCYL08, KMR^{+}05, KMS^{+}11, Kim05b, KW07b, KPK^{+}10, KI12, KMN94, KMMP12, KMS10, KMKP11, LM06a, LWDM12, LRG04, LFIM07, Li11a, LLM09, LABM^{+}12, LPA00, LLMR09, LGM11, LIBP07, MJYT06, MPV12, MOBS06, MEAC04, MM89, MTV09, MPT10, MZC06, MMO^{+}11a, MMO^{+}11b, MRSV07, MHA09, MMG10, MÁ05a, MS00c, MVA01, MAFSS04, MSY00, MBC^{+}97, NMM10, NSM07, NBGA08, NKC09, NK10a, NK10b, Ngu11a, Ngu11b, Ngu11c, NHJ12a, NHJ12b, Ngu12a, Ngu12b, Ngu12c, NL09, NV01b, OSR^{+}02, OBS12, OA07, ORB^{+}09, OBM11, OYA04, OT07, PWZ98, PCN10a, PCN10b, PCN10c, PdRM^{+}12, PDMFF12, PW09, PLC11, PA03b, PS11a, RA12, RL11a]. Intelligence [RCS10, RSTZ04, RTZŻ06, RTZZ08, RST^{+}10b, RST^{+}10a, RKS^{+}12a, RKS^{+}12b, SHN07, SÇS12, SZR^{+}08, SPCG07, SINK08, ShK06, Sav06, SDG^{+}07, STL11, Sco02, SKS^{+}09, SC09a, SCR06, SAG10a, SAG10b, SAUL11, SP07, SCB01, SKN05, TST10a, TST10b, TSCW11a, TSCW11b, TSJ12a, TSJ12b, TG04, TNN^{+}01, TN04, TNM05, TNVDF06, TNY07, TN08, TNI09, TND10, TNYL11, Tri98, TdRM^{+}09, VV03, VP04, WmCL07, WDGL10a, WDGL10b, WR11, WSN^{+}06, WSTI07, WY04, WZ08, WV99, YW06, YLC^{+}10, ZdC08, ZL97, ZL98, ZGY04, ZPTI09, ZO10, ZLY^{+}07a, Zui99, ZYW08, dRWK^{+}10, dCHM12, vdMS11, vM94, AZ96, APO96, AS98, AGS91, BBE01, BV05a, BSW95a, BA99, B^{+}12a, BL04a, BR92, BCD05, BJ86b, BFP04, BHR90, BK96, BHN97, BC04b, BCC99, CSCO00, CC93, CC95, CCP96, CP98a, CRL01, CHA03b, Coe98]. intelligence [Coe02, CS02a, CC97b, DDFM05, DdCF98, DBB^{+}04, Esp01, FD93, Fle02, FG96, Foo99, Fur05, GRT02, Giu98, GS95, GH96, Ham00, H^{+}05, HR91, HG98, HSL^{+}99, Kar91, KL05, KJ91, Ker97, KS93, Kow01, LM00, LM98b, Len97, MPP94, MPF95, MSZ90, MST92, McC96, MS02a, MSS03a, MN98, MHK05, MdPA98, Mit98, MS00b, NDF94, Nos88, NA97, Ohl93, OAdGBP00, PBB05, QBA01, Ral94, RM97a, RS99a, Reu97, Reu99, Reu01, RBW95, Sat97, Smi98, SM01, SV05, Tor93, VY93, VS02, WC95, Wei97, XCd03, ZL96, ZJ05, Zho01, dO98, dPMA98, AAW04, AN06, Cai11, CWZ^{+}11a, CWYC12, Che03a, CMZL04, CWY^{+}06, CYCL09, CChL^{+}10, KMR^{+}05, LIBP07, MZC06, MP01a, MBC^{+}97, OALZ^{+}08, Tri98, YZC^{+}07, YCC^{+}08, ZGK^{+}07, Zui99, dPMA98, Cad95]. Intelligent [ARSB09, AD06, ACK00, AMRŚ08, AHJ07a, AHJ07b, AHJ07c, APP12, BLB^{+}03, BSTL07, BTPPS06, BTPPS07, BTBP^{+}08, BTPPS09, BTBP^{+}10a, BTBP^{+}10b, BTKP^{+}11, BTPP^{+}12, BB10, Bou11, BKL^{+}12c, CFLR12, CHCA09, Coe11, CAB10, CKW11a, CKW11b, CSA^{+}12a, CSA^{+}12b, DGSM11a, DGSM11b, ERMS06, GHJ06a, GHJ06b, GHJ06c, GBH11, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, Gel04, Gel06, Gel07, Gel08, Gel09, Gel10, Gel11b, Gel11a, Gel12a, Gel12b, GZSC07, HRZK02, HKT12, HZH05, HLI06a, HLI06b, HLI06c, HHL07b, HHL07a, HWLJ08b, HWLJ08a, HJL^{+}09a, HJL^{+}09b, HZBF10, HZGZ10, HGBF12, HGGG12, HGPH12, Jac07b, Jac09a, Jac11c, JDAW12, KKGK11, KHJ05a, KHJ05b, KHJ05c, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KPRS07, KRSR11, LwCM00, LFJI10a, LFJI10b, LJS^{+}10, LHJ08a, LHJ08b, LHJ08c, MBW08, MMO^{+}11a, MMO^{+}11b, NHJ04a, NHJ04b, NHJ04c]. Intelligent [NBGA08, NM97, OA07, PWZ98, PPP07, PHJ03b, PHJ03a, Paz03, RO00a, RRBE09, RK12, SKS^{+}09, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, Ste09d, Stü09, VRHJ09a, VRHJ09b, Vis04, WPE05, ZS00, ZZ04b, ZZZL12, ZJL06, ZRTS03, BTPPS05, BMYZ95, CBN99, FKP^{+}05, Gel01, Gel02, Gel03, Gel05, Gom01, GLS05, HMRT05, Hag02, HS12b, HKB99, Hof01, Ima99, KHJ05d, LCB97, MVA01, Mul96, NcC99, RM97a, RZ91, RK93, RZ94, RM96, RS97b, RS99b, SYBM88, TTP05, AG98, Alb99, AKM10a, AKM10b, ABB^{+}10a, ACR^{+}08, ACD^{+}10, AAN00, BK05a, BGS^{+}03, BMN05, BHH08, CDCW09, CL01, CRW97, CGP02, CCPP12, CKO01, CYBF06, CY09, DG04, DFUR11, DAB01, FGM92, FGL96, FKLY08, FTC^{+}10, GHM05, GFV00, Goe98, GYA^{+}06, HLI06c, HJBF12, HMJG12, IAC06, JL00, JLS11a, JLS11b, JCC^{+}12, KBEP03, KL02]. Intelligent [LB03, LVP04, LwCM00, LCY03, LDXZ10a, LDXZ10b, LPA00, MSW05, MT02b, MA05c, MWJ97, MSR99, NNPW12, NLŚ10a, NLŚ10b, NKJ11a, NKJ11b, PCN12a, PCN12b, PCN12c, PGA^{+}05, PMA^{+}07, PLI08, RABR03, Rob93, RKNV09, SRW98, Smi06, SRL12a, SRL12b, SRL12c, VKMT11, WJ95, WMT96, WANL08, XXX^{+}09, XHXL08, XLHX08, YEY04, YAF^{+}02, YTC^{+}07, YWRS11, YCB12, YY01, ZZZL12]. Intensional [Urs02]. intensive [Bre97, WBHR08]. intentions [Sin94]. Inter [BB07]. Inter-Domain [BB07]. INTERACT [BPAB07a, BPAB07b, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CP05, EDGP11, GGK^{+}09a, GGK^{+}09b]. Interacting [Jac09b, SS11c, SS07b, SS11d, Rüt98]. Interaction [AJB09, BPAB07a, BPAB07b, BKLM07, BPR05, BD04, BSL^{+}09, BDLMM12, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CV04a, CGL^{+}09, CBG^{+}10, CBS^{+}12, Chi03, CFVCF^{+}12, CKOB11, DGSM11a, DGSM11b, DGWB09, EDGP11, EKF12, EBAH08, FdCLN^{+}11, GCK06, GGK^{+}09a, GGK^{+}09b, HHP09, Hol07, Hol08, HM09, HS11, HSL06, Jac07c, Jac07d, Jac07b, Jac07a, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, KMPH09, KW10, LK09a, LT11a, LCHS08, LHH10, LSHB07, LN01b, MSB12, MB06, NSFB10, OB07, PPP07, PB08a, PB08b, PBRB08, PBS08, RBF06, SS09a, SLH04, SMF^{+}11, SS04, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, SS06e, SKM07, vDLJ11, BUG93, BB05a, BR06, BGU94, BGU95, Bra99, BMS01, CP05, FKKB^{+}05, GT90, GT93, Ish99a]. interaction [MJR04, Pat02, SLH05, TTP05, TC95, WF98, WS02, Ish10]. interactions [Pai00]. Interactive [ADM^{+}06, ADM^{+}08, ALL^{+}10, BPR05, BMO02, BDMS08, CFVV^{+}12, FdCLJ^{+}11, FP08, GSH^{+}04, GMI06, GHP^{+}08, IZP09, RR04b, STA^{+}11, SS08b, VC03, BL05a, LSW81, MSW05, ALL^{+}10, BC89, BF12, BH77, BMP96, CCJ07, DOS99, DB07, FLUV02, GH06, GP98, GP08b, Her02, IZP09, Joh01, JJF03, KP10a, PP01, RR04b, SvS00, She01, STA^{+}11, SS08b, SW97, Vaz99, VC03, ZPT^{+}06, vEGSW11]. Interchange [BGP08, GHP09, PB07, ZZZL12]. Interconnection [Sch98c]. Intercultural [IFV07]. Interdisciplinary [Akm01, BCRB^{+}11, Bla03a, BSB^{+}99, DKLT05, FHW98, KRRBV07, PR07, SW75]. Interest [BS10]. Interface [CHD07, CLS07a, CDND11, DKP00, DLO03, EPVW10, KGRD10, KKDV02, KKD04, LKD08, MTWD06, RWD09, SS09a, SS11c, SS07c, SS07b, SS09c, SS11d, TBD12, WJP07, AD98, BDW97, CD01, DKD05, DLM99, HR99a]. Interfaces [Ayk07a, BC02, CS03, CFVV^{+}12, ECV^{+}10, EEM^{+}11, FdCLJ^{+}11, SS04, SP07, TSG00, Wal84, BS76, Bee86, DEM81, GT05, Mit98, PCA02, IM10, Pai00]. interior [Koj91]. Interlaken [CC04, DGN01]. interlingua [EEM96]. intermediate [Goo83]. Intermittent [DLdS04]. International [ÅBE12, AH00, AO02, AAW04, AG01, AI97, AD05, APFV09, AB10, ARS03, ARS09, AK07, AP09, AL11a, AP93, AGB^{+}11, AKZ01, Abe10, ABB12, AKK04, APR^{+}06, ACN^{+}07, AK90, AM10, AJNJ12, ABS01, AMS06, Abr07, AM91b, AM91a, Abr01, AGK^{+}10, ABD^{+}09, AS11a, AI06, ADG^{+}08b, AHS11a, AHS11b, AB11, ABG^{+}06, ASS10, AJRB09, ABH12, ARSB09, AST05, AN97, AFvZ02, AGP08, AK97, AR99, ABB^{+}11, AAA^{+}11, AFP^{+}10, AS99, ADDL08, AW12, AYV^{+}11, Agu12, ABG09a, AW01, AF97, AM03, AdBB12, AB12, APS08, ABKT11, Akl95, ACD^{+}07, Akm01, ABCWM99, AMU03, ASE^{+}07, ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12, ASP01, AG98, Alb99, AMSM^{+}09b, AJM87, AS03, ADP09, ABDW10, AKM10a, AKM10b, ADJ^{+}01a, ADJ^{+}01b, AvAS06, AvASD06a, AvASD06b, AvASD06c, ABMS06, AGL^{+}07]. International [AE05, AD06, APPE09a, APPE09b, AV07, ABB^{+}10a, AF11, ACW07, AAB^{+}06, AH02, Alo05, ADR07, ABB10b, APSZ02, AV08, ACJ09, Alp11, ANR99, AM07, ABB99, AFM01, AJB09, AS11b, ANF08, ABR09, AR10, AL03a, AP04b, AP04a, APBP07, AL03b, Ama08a, Ama08b, AC11, ACK00, AYBH^{+}06, ABD^{+}99, Ami98, ATM09, ASR^{+}07, AMRŚ08, ALPH10, AFG^{+}11, AFS02, AHK07, ABS91, AFGvE10, AFGvE11, And96, ACH00, ABM02, AFL03, APW06, ABJ02, AH98, ABH^{+}10, ADM^{+}06, ACL98, AMCL02, AYB^{+}09, ABK12, ATO^{+}09, APHA11, ALR11, AIL12, AB04, AAB^{+}07, Ant97, AKL^{+}99, AT07, APN08, AT12, AJ11, ABG96, AMNT06, AHJ07a, AHJ07b, AHJ07c, AGI^{+}06, AFR^{+}04, AGM03, AT02c, AS07a, AS10, AÖ12, AS12, ABGL09, AZ11, ABM05, AS07b, AJ94, AM98, AF99b, AM00, AJS00, Arm02, ABD08]. International [AJRS03, AWA^{+}11a, AWA^{+}11b, AN07, AN09, AOSV08, APV09, AM02, AKG06, AP07, AHMT08, Asa99, AKR03, Asa06, AiNOW11, AB06, AB03, APP12, ABD03, ABT04, Asp98, ASAM07, ABW10, ABRZ06, AH10, ALZ11, AJ01a, AJ01b, AMM99, ACR12, AHSC11, ACHH12, Auc98, APM08, ADP11, Aus86, ACR^{+}08, ACD^{+}10, AKTT12, AKS09, Avg12, AKSO06, AOM07a, AOM07b, ADS09, ADGM12a, ADGM12b, ADGM12c, AGCP^{+}11, ABRS06, ADK04, Ayk07a, Ayk07b, Ayk09, AS07c, ABFR07, AKR^{+}09, Baa03, Baa07, BSMM04, BV02, BM03a, BM96, BVO10, BBMB^{+}08, BBD^{+}10, BW05a, Bac00, BO00a, BGHJ07, BC09a, BBCL08, BG91a, BM99a, BBdB12, BS11a, BA06, BW03a, BH06a, BMS^{+}08a, BCBF08, Bak95, BBT08, BHL^{+}03, BMK08, BMK09, BDK10a, BKH11, BC95, BLS06, BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, BST01, BST03, BR03]. International [BJ06, BdBC01, BSC12, BS09, BJ10, BFGM05, BTT90, BB07, BCT02a, BMUV10, Ban92, Ban93, Ban94, BKS99, BR12a, BDZ04, BLO^{+}07, BLW09, BYLJ11, BW11, BSZ12, BGG^{+}12, BBF^{+}06, BP04a, BGLT05, BBS07, BPAB07a, BPAB07b, BKA11, Bar08a, BBB^{+}07a, BL12, BP12a, BN94, BH06b, BFH07, BCS09, BHH04b, BHH04c, BCC06, BPCS06, BKLM07, BK05a, BHKR11, BBH^{+}10, BBA12, BHJKPG06, BBM^{+}08, BG93a, BFF^{+}10, BGK12, Bar03, BDPS02, BGH06a, BdB08, BH10, BPS11, BLS12, BDE12, BG09, BBAB^{+}07, BL06a, BRS^{+}07, BW03b, BR04, BGP08, BGP11, Bat01, BS11d, BS11e, Bat09, BW10a, BGV01, BMP00, BMP01, BTQ02, BTQ03, BBP04, BPW06, Bau99, BHP95, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BBP^{+}12a, BBP^{+}12b]. International [BD12, BS06a, BPV04, BPR08, BWAM09, BH08, BM11a, BDG12a, BBK12, BBDS12, BB99, BS06b, BL05b, BCRB^{+}11, BB95, BC92, BdQ11, BJP12, BDIR07a, BDIR07b, BPR02, Bel03, BMR^{+}04, BL08, BHRU09, Bel00, BP11, BCC^{+}10, BR92, BE08, BC10a, BGKY08, BEP04, BBB07b, BDCM04, BS03a, BCG^{+}07, BCKR10, BCST99, BKR09, BDG^{+}12b, BG11a, Ben06, BP09a, BP03, BV92, BLJ08, BC96a, BCD04, BDd09, BDW08, BMSHW09, BFBS11, BF09, Ber10, BS92, BMS04a, BMS08b, BJM09, BNUW09, BF12, BR89, BCG^{+}11, BI03, BH07, BDZ08, BPZ09, BI11, BCP12, BFGL10, BKH^{+}09, BL10a, BC11a, BF11, BC06, Ber98, BCM00, Ber02b, BBKW03, BCV06, BDSW09, BLB^{+}03, BGF06, BSTL07, BU94, BCP^{+}04, BRL08, Ber03, BDH^{+}99, BR97, BCP99, Bes07]. International [Bes93, BKL^{+}12a, BKL^{+}12b, BG11b, BB01b, BND01, BG93b, BLSS11, BM99b, Bha00, Bha02, Bha05, Bha07, BKW08, BKW09a, BW10b, BB03, BM10, BD97b, BN04, BG06, BN91a, BK05b, BVAT12, BCB97, BS01a, Bih97, Bih03, BY07, BG12b, Bil97, BdBK03, BR96a, BZU07, Bir12, BMS95, BGS11, BV11, BH92, BBKM11, BBRM98, BHL90, BBP96, BBZ00, BBZ01, BSS11, BV12, BSBM05, Bla03a, BvDMST11, BMS04b, BTPPS06, BTPPS07, BTBP^{+}08, BTPPS09, BTBP^{+}10a, BTBP^{+}10b, BTKP^{+}11, BTPP^{+}12, BBH01, Bla90, Bla03b, BMSW12, BBRS05, BBC^{+}08, BBD^{+}09, BBR^{+}10, BS02b, BJN03, BPT06, BRST08, BK02, BB10, BKV10, BMO02, BPS06, BHHJ11, BRW93, BLKW01, BR12b, Bod05, Bod03, BL06b, BRU06a, BJS99, BGPW05, BKLP11, BKLR12, BHU03, BFSN10, BPD06, BCC^{+}12, BM02a, BDS04, BMN05, BKW09b]. International [BTCW10a, BTCW10b, BMW10, BTCW12, BG10b, BS94, BJ01a, BDK10b, BTZ^{+}10, BO00b, BKS01, BI04, BKS05, BLvdH10, BJ12a, BB06, BFMS08, BFJM09, Bon03, BDDM12, BJ07, Boo91, BF89, BDDS05, BDDS06, BDDS07, BNS00, BGR03, BBBB08, BG07a, BG08, Bor94, BGPR10, BH02, BF06, BS93, Bos01, BK04a, BGL09, BL10b, BM02b, BHJ^{+}11, Bos00a, BL02a, BE07, BGJ11, Bos00b, Bos96, BR07a, BF10, BM09a, BC10c, Bou11, BM94, BMYZ91, BMYZ95, BMCCM11, BCRT92, BFMR96a, BFMR96b, BKM08a, BHL11, BDN08, BCK10, BJ01b, BIMW12, BSB^{+}99, BZ08a, BP09b, BEM06, BAP^{+}05, BKL^{+}12c, BKM01, BGKS04, BBPP09, BH95, BHT97, BFH98, BDGK00, BDGN07, Boy01, BM03c, BC11b, BKLS11, BR07b, BR09, BS01b, BL07b, BDV95, Bra02, Bra99, BRT11, BTT12, Bra09b, BW00, BC11c]. International [BL01, BKM07, BLV02, BAH09, BBT10, BCF06, BH00, BKZ05, BNZ06, BZ09, BB11, BHV11, BW01, BC09b, BHH^{+}06, BFvdM10, BGHS05, BBK^{+}11, BCH^{+}05, BKW06, BKNS06, BMS01, BD04, Bri93, BAdVR08, BJKK02, BHLvdP07, BBH08, BHQ10a, BHQ10b, BUHS08, BGR08, Bri95, Bri97, BL02b, BRP09, BFMS01, BEFP08, BP12b, BR98, BZ03, BJ12b, BtHRBB12, BSH06, BHJY07, Bru06b, BW09, BD11, BS12, BK04b, BM11b, BSS00, BCR03, BW92, BP93, Bru04, BSL^{+}09, BHM03, B^{+}12b, BAWH00, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BvADS08a, BvADS08b, BMC08, BV94, BM06, BCF^{+}02, BD08, BS01c, BBL87, Bud91, Bue04, BPSW06a, BPSW06b, Buh98, BBB06, BHH08, BOH^{+}12, BH11, BLPW02, Bun94, Bun96, BZ08b, BKB03, BC98, BE01, BD03, BR07c, BLVV07, Bur90, BY06, BTMI11]. International [BL06d, BDLMM12, BHM^{+}02, BkHJ^{+}04, BFOY04, BH06d, BCDN07, BLS02, BRS09, BJ01d, BPS02, BHLP07, BS11f, BGH06b, But10, BKO03, BKO04, BFKO06, BFK^{+}07, BFK^{+}08, BFC^{+}09, BB00b, CPS05, CSPC09, CRJ11a, CRJ11b, CV11, CC04, CAD^{+}02, CCL06, CLKL09, CHKL10, CV07, CSCO00, COBV11, CD10, CG07, CBGCCZ12, CYM^{+}11, CLMP02, CC93, CC95, CL96, CP98a, CBC^{+}02, CIW06, CDP02, CDV03, CCD04, CDP^{+}06, CCF^{+}08, CCD^{+}09, CHM^{+}10, CKPR11, CLM03, CL08a, CKS04, CPR^{+}10, CKP^{+}12, CCJS07, CKD11, CK12a, CRL01, CK04a, CK04b, CK06b, CK06a, CK08, CK10a, CK10b, CK12b, CK12c, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CV04a, CV04b, Can12, CSJD06, CLWC08, CGL^{+}09, CBG^{+}10, CFZ10, CZF10, CBS^{+}12, CHB^{+}12, CV08, CMSM^{+}10, CP12, CS03, CS11, CS05a, CDCW09, CZ02, CS07, CP10a]. International [COT98, CNP^{+}08, CMT02, CBF09, CP10b, CT06, CFM09, CF00a, CPSA10, CGMP11, CGS01, CVTP12, CJ08a, CL01, CFGN11, CFH^{+}12, CDGW10, CD07, CA08, CL07, CR01a, Cer05, CP11, CC07, CD00a, CGP02, CCPP12, CVV08, CBNR02, aCF06, CW11, CCK^{+}08, CY03, CT07, CSM10, Cha05a, CM12, CCD^{+}06a, CCD^{+}06b, CMZ99, CM03a, CG02, CGBZ08, Cha02a, CCL02, CWC^{+}07, CKK^{+}09, CHCM11, Cha05b, CZG^{+}08, CGKM11, CMMR12, CJRR07, CC91, Cha02b, CPC^{+}99, CH10, CDPT06, CMM^{+}09a, CSR08, CL10, CWiM^{+}09, CW09, CLL^{+}09, CGL^{+}11, CJRT05, Che03b, CR04a, CCM^{+}04, CH04, CL06b, CWY^{+}06, CDLZ06, CPZ07, CJ08b, CRW08, CMS08a, CLZ^{+}09, CLLD09, CF09a, CMM09b, Che10, CY10, CTYG10, CTS10, CWZ11b, Che11, CEK11, CY11a, CFLR12, CC12, CYZ12, CP06b, CLZ06, CEQ^{+}11, CCP10a, CCP10b, CHG^{+}07, CNA08, CCdS03]. International [CFK^{+}12, CD00b, CvSW04, CM03b, CHCA09, CY12, Chi95, CHZ08, Chi03, CBS^{+}11, CC97a, CH09, CKO01, Cho02a, Cho02b, CK06c, Cho12, CQSS09, CW00, CRD08, CKH10, CCAS12, CCBW11, CH99, CBN99, CS05b, CHAL04, CDY^{+}11, CCLR98, Chr99, Chr00a, Chr01, CCMR02, CCMR03, CCMR05, CCMR06, CCMR07, CCMR09, CCMR10, CMMR11, CCMS11, CMSA12, Chr91, Chr00b, CKM^{+}09, CHA03b, Chu03, CM06, CSY09, CY11b, CI98, CM04a, CH96, CW99, CW01, CW06, CM09b, CGP03, CB06, CS12a, CP10c, CP99, CFVV^{+}12, CFVCF^{+}12, CHD98, CLB11, CPPB06, CCM08, CCK95, CA10, CV10, CVV12, CER79, CG09, CU08, Cle92, Cle99, CGS07, CS00b, CH98, Cod01, Coe11, CGK^{+}12, CCD^{+}12a, CCD^{+}12b, Coe02, CPvdH03, CSTD08, CF09b, CBT07, CMM93]. International [CGM95, CAB10, Coh10, CMS11, CÓ11, Coi99, CH91, Col02, CML^{+}10, CDN12, CK01, CGS02, Com97, CBF01, CMT01, CMP07, CDSS07, CR01b, CLS07a, CM00, CDW86, CMV96, CNmWH^{+}05, CGGO03, CS02b, CP07, CKOB11, CMCS11, Cop95, CDNS00, CYBF06, CAP08, CWO^{+}09, CY09, CRS10, CKW11a, CKW11b, CSA^{+}12a, CSA^{+}12b, CAdB01, CS97, CFP^{+}09, CBW99, CEKR02, CEM^{+}06, CM09c, CKC11, CTS07, CR04b, CF99, Cor02, CCSW12, CDFP11, CM04b, CSRS08, Cou93b, CS96, Cou01, Cou03, Cou05, CM11, CMB^{+}08, Cov99, CPR12, CS01b, Cra05, Cra08, CL08b, CS12c, CvRVSN12, CBBS09, CRYS09, CMOZM11, CPP07, CMM01, CDM04, CR05b, CFS06, CSSW04, CFL10, CRW^{+}12, CFRR10, CRS00, CH01, Cro03, CDT11, CDA^{+}06, CK93, CDG89, CVÉ07, CGBDA08, CMS08b, CLBP11, COB06, Cue03]. International [CMQ96, CEER96, Cur09, CFB11, CF00b, CKLM09, CD11, CD12, CT99, CE00, COB^{+}08, CMPW12a, CMPW12b, DGT10a, DGT10b, DGSM11a, DGSM11b, DFL^{+}10, DR02a, DW03a, DN07, Dai07, Dal92, DKY06, DH05, DYCD09, DO02, DH07a, DL06b, DGT10c, DLV04, DM06, DG06a, Dan12, DF10, DS05a, DD97, DGT96, DB02, DD03, DSDS04, DG04, DPLW08, DU09, DK90, DDE04, DSLT08, DSRW08, DSLT10, DSHL11, DH01, DG06b, DSZ^{+}11, DFUR11, DS09a, DMO98, DF01b, DFR02, DSM94, DM01b, DSB04, DMS04, DJAP07, DG07a, DCFR12, DV10, DdQ10, DW07a, DEKR09, DDHP05, DPC^{+}07, DG01, DAB01, DKS^{+}07, DKC10, DSB10, DLNU11, DC12, DH97, DMR01, DF09, De 07a, DR11, DD95, DFM04, DY06a, De 10a, De 10b, DKK08, DFPV11, DW12, DG02, DGJ^{+}05, DHRT10, DZ10a, DS09b, DBC^{+}10, DMW00, DRDKE11]. International [Dec00, DGG^{+}11a, DIMV09, DFMV10, DMV12, DPV11, DG09a, DGM09, DG10, DEG11, DG12, DFK91, Deh97, Deh99, DST01, DSS03, DSZ07, DGST09, DIS11, DKM^{+}05, Del00, DF11, DFP^{+}12, DKM91, DDJ00, DS01a, DP11, Dem07, DQBZ02, DBPZ05, DY06b, DG07b, DWWL09, DMLW11a, DMLW11b, DMLW11c, DBB12a, DV07a, DFG^{+}12, DGLT12, Der89, DL95a, DV07b, DLW12, DS98, DPW04, DH10a, Des02, Des09, DJS06, DJNR06, DLN10, DGSN11, DKN^{+}11, DSM03, DKLT05, DCP95, Di 97, DSV01, DR06, DPP09, DPT12, DR09, DB09, DP10a, DDI10a, DDI10b, DDI10c, DGWB09, DO89a, DO89b, DOS99, DKLS04, DJRZ06, DPA07, DD10, DVM^{+}11, DVV07, DN09, DC00, DMWW88, DJJ12, DD07, Die03, Dig04, DvEF06, DTMH10, DGR06, DR02b, DGDM^{+}11, DMRS06, DMRS07, DZGS06, Dim03, DDK11, DSZ10]. International [DS03, Din03, DG09b, DLdS04, DMB^{+}07, DKN12, DJNR09, DKU05, Dit88, DGN01, Dit01, DFN97, DPP97, DPP98, DL04b, DFN10, DLGJ10, Dji89, DLŠ11a, DLŠ11b, DPT08, DB07, DK02a, DK02b, DSL08, Dol06, DHO08, Dol09, DO09, DCFY10, DO11, DS11c, DIP12, DFT04, DFF06, DFPS06, DFS08, DFM11, DFT12, DPMP^{+}11a, DPMP^{+}11b, DPSU11, DP12, DT06a, DGG^{+}10, DTW03, DW03b, DLW^{+}07, DDI09, DZ10b, DW94, DMW96, DP97, DMW06, DPNM12, DBH01, DDS02, DGB^{+}06, DBB^{+}08b, DBD^{+}10b, DBB^{+}12b, Dor02, DGJ09, DF06, DHMM96, DFD04, DW07b, DXZ07, DTF98, DKR02, DZ94, Du00, DHP09, DSS^{+}07, DP06, Dub09, DKOZ05b, DKOZ05a, Due04, Duf07, Duf09, Duf11, DHH^{+}08, DRS08, DH08, DA01, DBB^{+}08c, DG97, DKN02, Dun09, DS06, DH07b, Dur12, DLM07, DLJ12, DFGN12]. International [DFS06, Dwa00, Dwo06, Dwy01, DL07, DNS95, Dyc00, DF99, DS07, ET01a, EV96, EKS10, EB04, EM03, ED06a, EBT11, ELS12, EK09, EKF12, EM02, EGMW11, EM08, EKO08, EGO09, EF02, Ehr98, EFG^{+}09, ENR83, Ehr85a, Ehr85b, Ehr87, EEK^{+}87, EKLM87, EKR91, EEKR00, EHRT08, ERRS10, EEKR12, EBSS98, EV01, EdS08, ES10, EM05, EFT01, ES02, EK12, EEM96, Ekl04, EH08, EKR10, ECB06, EKP^{+}07, EHH12, EM06, EK10, Ell95, EP03a, EP03b, EKT94, ELNM08, ELN^{+}10, EML^{+}12, EHM11, EG97, EOR06, ES00a, EN06, ET01b, EW04, ES03, EOSW07, EPVW10, EBT07, EKO^{+}04, EG10, ELS09, ÉEMS04, EW03, EL02a, EP06, EK07, ENO12, EWZ11, Esc10, EV07, EF03, Ési06, EL02b, EM10, ERMS06, EFZKM07, EBAH08, EV09, EHMM09, ECV^{+}10, EVV^{+}11, EEM^{+}11]. International [Est98, Est99, ET06, ES92, ES00b, EKM06, ED06b, EKS00, EH12, FRvKS10, FHJ^{+}02, FRS08, FT11, FSA^{+}07, FBKS03, FBSMS08, FL00, FV01, FV03, FV04, FV05, FPFR05, FWY05, FPP11, FZJE^{+}05, FZJE^{+}06, FD03, FMC04, Fed01, Feh11, FR01a, FRS02, FRS03, Fek08, FR11, FKK09, FKP99, FWZH06, FSM03, FKB02, FV10, FLVS11, FSdlPT11a, FSdlPT11b, Fer08, FR09, FR95b, Fer96, Fer98, FMZ05, FIAP01, FFL^{+}12, FM04, Fes10, FL09, Fia99, FS07c, FI08, FKK04, FKM09, FdCLJ^{+}11, FdCLN^{+}11, Fic03, FMP11a, FMP11b, FMP11c, FMW01, FK02, Fie02, FV09, FOGE^{+}09, FZJ01, FPCK11, Fin99, FNY93, FLP12a, FHFL06, FHLP09, FHH11, FHW12, FHKQ12, FDL10, FW07, FBM12, Fis08, FST09, FvdTDG12, FSN06, Fit84, Fit93, FJL97, FHT05, FHY08]. International [FBV11, FK12a, FB09, FX08, Fle99, FG06, FLF^{+}08, FKS10, FR99a, FSV09, Fom06, FCLL02, FKW08, Fon03, FRL07, FG96, FLUV02, FP08, FGK^{+}11, FMdGC99, Fos99, FHP03, FILP06, FNPW05, FS07d, FS08, FS09b, FS10, Fra00, FR99b, FKBA12, FW09, FRSH05, Fra01, Fra99, Fra04, FHW08, FGK^{+}10, FL11, FG10, FdS12, FGM92, FGL96, FSTS09, FCD^{+}04, FLL03, FKM08, Fre98, Fre01, FM99, FNGW08, Fre96, FPRS06, FGTZ08, FLR11, FT94, Fri01, FMRW09, FHP11, FSP09, Fri93, Fri94, FD11, Fuc98, FLMK06, FLT07, FKLT08, FvdHJT10, FK09, Fum97, FS06, FN12b, FKL08, FLP11, FCDB07, FM96, FMY04, FKLY08, FTC^{+}10, GO96, Gab97, GVO^{+}10, Gab13, GHJ06a, GHJ06b, GHJ06c, Gac02, Gae96, GGD09, GP07, GP09a, GP11, GPAS^{+}06, Gal07, GP08a, GN12a, GTT04]. International [Gal97a, Gal97b, GCJB09, GBH11, GLP12, GE08, GM00, GMV06, GMV07, Gan99, GVM99, GWY99, GML^{+}08a, GLSY10, GPCL10, GLW^{+}12, GH03, GLMP07, GMO09, GD10, GANACBR10, GANACL11, GANACBdCdV12, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GVCA07, GMS03, GY04, GG09, GWK09, GH08, GY08, GM01, GS01, GLV09, GVT08, GL10a, GJ93, GW96, GNR07, Gau93, GFV00, GJT03, GLZZ09, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, GGK06a, GHR89, GLCT10, GKC^{+}12, Gec81, GMV03, Gel01, Gel02, Gel03, Gel04, GdATM05, Gel06, GRG06, Gel07, GM07a, Gel08, GM08a, Gel09, Gel10, Gel11b, Gel11a, Gel12a, Gel12b, GLP99, Gel99, Gen09, GH94a, GH94b, GRB^{+}07, GM03, GLW07, GB97, GFV09, GKS^{+}07, GMV09a, GKMV10, GKMV11, GKMV12, GK06, GGHN97, GL10b, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, GG07a]. International [GG07b, GG07c, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, GMP06, GW03, GKT09, GKT10, GKST11, GKS12, GG02a, GHP^{+}11, GPR^{+}12, GS09a, GGS09, GDP07, GH07a, GO11, GM12, Gia89, GAA06, GO09, Gib12, GN12c, GCK06, GW09, Gie10, GKL^{+}11, GR12, Gie05, GH10, GMBM05, GH05, Gil10, GG98, GSYZ10, GS09b, GH12a, GH06, GHI^{+}12, GPM98, GMO03, GHSW04, Gir99, Giu98, GOW03, GT04, GHK97, GOZ05, GGS11, GZR02, GH09, GL74a, GL74b, GSH^{+}04, GMI06, GMUW12, GM06a, God05a, GP09b, GP98, GJRR08, GSS09, GJN10, GJRT01, Goe98, GK01, GW11a, GCSR07, GZ09, GA09, GJRR11, Gol96, GLIC10, GF12a, GGHS07, GSLM12, GPB02, Gom01, GHSY06, GG10, GLC^{+}11a, GLC^{+}11b, GdlP99, GCM08, GP12, GM99, GK02, GW98, GQ11, Gor03, GLN01, GSP01, GPS03]. International [GLS05, GKS05, GZSC07, GW06a, Gór06, GHC^{+}06, GH12b, GS88, GN90, GV95, GGS99, GR06b, GM06b, GHP09, GS10, GN05, GBMS11, GS00b, GM04b, GZ06, GP08b, GMS12, GSW06, GYA^{+}06, GHH^{+}06a, GL08, GCC^{+}02b, GTY03, GSS00, GW06b, GL03, GHH06b, GS08, GM11b, GN08, GGK^{+}09a, GGK^{+}09b, GSS11, GW12, GF12b, GH04, GH07b, GH93, GN98, GRP10, GDGB12, GMV12, GAR03, GAR04, Gue05, GRR05, GS06, GB07, Gue01, Gue04, GP09c, GG02b, GQ95, GW11b, GJTS06, GF97, Gun94, GY03, GYD^{+}06, Gup99, GM08b, Gup10, GM10, GJRS12, GKOT00, GM07b, GPL99, GDG11, Gyi96, GPV88, HNH^{+}09, HP02a, HOC07, HM11a, HRZK02, HP12, Had03, Hae99, HM11b, Hag02, HM94, HO03, HW06a, HKR07, Hal94, HP99a, HENR03, Hal09, HG02, HS12b, HPTV05, HT97]. International [HCT02, Ham08, HT09, HMT10a, HLSZ11a, HLSZ11b, HT11, HHMT12, Ham91, HS07, HDC99, HTW02, HY12, HP99b, HV03, HMS09, HWW^{+}10, HKB99, HMT10b, Han07, Han09, HLC^{+}12, HS06, HB97, HAM95, HKP08, HKL07, HHLS08, Har01, HRC06, Har07, Har09, Har11a, HS08a, HMTH10, HP95, HS94, HG96, HK98a, HG00, HKI08, HZK08, HH10, HMT99, HZ10, HS08b, HK92, HPV00, HNRW06, HMP08, HBS00, HM12b, HKS06, HHP09, HLKX12, Heb95, Hed03, Hee94, HNM02, HKP10, HU08, HLW04, HMW^{+}08, HR95, HJ12, HJMP06, Hen08, HA02, HWG10, HK10a, HM03, Hen06, HS98a, HK01, HBN06, HRB12, Her00, HV06, HS02, HP06, HP94, HS95a, HP02b, HC11, HMM12, Her92b, HAB98, HHJZ12, HHLR08, HS95b, HHW01, HT06a, HPP06, Heu98]. International [Heu00, HN12, HA00, Hic04, Hig05, HSS07, Hil06, HW09, HRT^{+}03, HRTR05, HRR^{+}06, HPS09, HRSS08, HKLP12, Hir97, Hir98, HF97, HL10, HS95c, HK10b, HCF06, HZ08, HSKB09, HGZ05, HJRS99, HAB02b, HW07, HdQ08, Hof01, Hof03b, Hof11, HCR06, HW03, HK96, HKT12, HSB^{+}10, HŽ07, HB11, HKP11, Hon01, HNF08, HNQ08, HI10, HDGK11a, HDGK11b, HM01, HPRS06, HM05a, HRM^{+}10, HMMW10, HVT12, HRRY08, HNP12, Hor00, HMG03, HSH08, HCDL09, HH02, Hor02, HY03, HZB07, HMPZ09, HPP88, HS98c, HNW04, HKV10, HL91, HK98b, HYPY10a, HYPY10b, HYMZ11, HV98, HRA02, HW08a, HLCZ11, HSSV12, HM12c, HdL12, HC09, HSB^{+}96, HH03, HZH05, HLI06a, HLI06b, HLI06c, HHL07b, HHL07a, HNPP07, HKM^{+}07, HWLJ08b, HWLJ08a, HYP^{+}09, HJL^{+}09a, HJL^{+}09b, HZBF10, HZGZ10, HGBF12, HGGG12, HGPH12, HJBF12, HMJG12]. International [HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, HMBD09, HW08b, HSMPA09, HKS09, HP89, HL00, cHPkC^{+}07, HS99, HMT10c, HKH10, HLFS12, HS08c, HHG11, HJ00, HS03, HMCL06, Hus01, Hut94, Hut96, HK07, HDLM09, HSTU99, HMSU04, HU05, HST07, HSVZ10, HdMKS12, IdKM06, Iba92, IKO03, IZ02, ID03, ID06, IY06, IR08, IKR00, ICSY07, IT06, IPSK04, ILR95, IMW04, IMK06, IAC06, IS11a, IS11b, IRM93, IZ98, IZ99, IZ00, Ima99, IR07, IMY^{+}07, IPRO08, IST07, IL01, IJ06, IFGH10, IFL^{+}12, IKY05, ISS^{+}06, IS98, IK12, Ish99b, IGN05, IFV07, Ish97, IHX^{+}08, IDMY08a, IDMY08b, IS02, IS12a, IS12b, IS08, IM91, IY95, IHQ96, IT03, IKKU08, IT08, IZP09, IS12c, IN11, JZR09, Jac07c, Jac07d, Jac07b, Jac07a, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c]. International [Jac11d, Jac04, JRSN09, Jaf99, JMBS07, Jäh99, JSV11, JOD98, JM05, JM11a, JYZ04, JRL09, JM07, JM10, JR98, JK00, JLV02, JSO04, JKRR04, JSO11, Jan88, JSWhT11, Jan89, Jan92, JS01, JLST06, JT09, JBJ94, JO99, Jay99, Jay04, JLM00, JFBH08, JNR11a, JNR11b, JZ03, JS08, JNHJ10a, JNHJ10b, JNH11a, JNH11b, JSW03, JL00, Jen01, JJP^{+}02, Jen03, JP04a, JPD04, JLS11a, JLS11b, JM95, Jeu98, JCC^{+}12, JDL11, JHC02, JKN^{+}12, JSK10, Jha10, JM11b, JS11a, JWH06, JP09a, JNPV10a, JNPV10b, JNPV10c, JFT11, JDAW12, JWbG^{+}06, JWG^{+}06, JPXS04a, JPXS04b, JGXC04, JRPP07, Joh03, JM03, Joh01, JV06, JP11a, JK04, Jón12a, Jón12b, JLW07b, JMO09, Jon00, Jon01, JS02, JP96, JVA04, JJF03, Jos00, JDMB08, JDMB10, JMP12, JS11b, Jou09, Jou11, JY14, JS94]. International [JQ04, JMO10, JMT11, Jue04, JP12b, JK06, JY12, JRR03, JN12, Kåg98, KEDW07, KS03a, KSG^{+}09, KG12, Kah79, KMP84, Kah03, KG04, KS08a, KN09a, KPS12, KM11, KKP02, KCR04, KMT00, KWA01, KWA02, KMW03, KMW04, KC07, KC09, KKGK11, KC11a, KC11b, KS04a, KM08a, KJR05, KK97, KLZ07a, KLZ07b, KCYL08, KR10a, KA12, KPSW10, KMR^{+}05, KL07a, KSS08, KO91, KMMN11, KvETvdH10b, KvETvdH10a, Kap92, KR02a, KR04a, KJ09, KTH09, KS03b, KMD02, Kar77, Kar83, KV04, Kar09, KL07b, KP92, KT96, KMC^{+}01, Kat99, KS02c, KK11a, KK11b, KLP04, KLB^{+}06, KLS10, Kat92, KY07a, KS09b, KWC^{+}12, KKMW07, KW07a, KP10a, KLT^{+}12, Kay03, KP04, KLM^{+}07, Kei09, KJ91, KMF06, KS03c, Ker01, KBP07, Ker12, KMS^{+}11, KLP10, KS12b, KL12, KSB10]. International [KMSB11, Kil01, KM96, Kim01b, Kim01a, Kim02, Kim05b, KS06, KYL07, KKH09, hKhLKŚ10, KSB11, hKAS^{+}11, Kim12, KM04, KWCW06a, KWCW06b, KWCW06c, Kin08, KjHGG11, KDS01, Kir93, KK98, KR00a, KR02b, KJB09, KILW10a, KILW10b, KR00b, KR01, KN03, KSUZ11, Kla12, KD99, KY07b, KHG01, KPS01, KŽ04, Klu97, KW98, KSW99, KK00, KZ01, KOS02, KOOL03, KOK04, KRP06, KHPS07, KPP08, Kno11, Knu99a, Knu02, KLLP12, KP01a, KP99, KP01b, KNP01, KFW04, KH11a, KH11b, KKM^{+}11, KH02, KMSS02, KRRBV07, KTB09, KHL10, KSDO06b, KSDO06a, KK11d, KBHSW06, KHSM^{+}08, KK11e, KS05a, KW07c, KHW03, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KCMO11, KLPP03, KC02, KKS10, KC00, KPS09, KvdM00, KFMM06, KW10, KKC09a, KKC09b, KV08, KK09b, KBH03, Kos98, KY11, KKMN07, KBČ^{+}12, KS10a, KS12c, KMS12, KM00]. International [KYT01, KHI06, KHS07, KMMT12, KGS08, KU12, KLT^{+}07, KHK^{+}08, KVMH09, KP09, KR10b, Koz04, KH11c, Kra03, KU06, KN09b, KP05, KLN07, KO07a, KKL12, KT11, Kra99, Kra98, KJ10, KSHM09, KR02c, KO07b, KP10b, KPR07, KP11, KP15, KOS07, KMVY06, KMYJ08, KKH07, Kro00, KS12d, KLC07, KASS07, Kru08, KNK08a, KNK08b, KPRS07, KRSR11, KSTP99, KNP06, Kuc̆02, KK07, KP10c, KS96, KU01, Kuc11, KMN94, KWT03, KF12, KRS02, KR04b, KH12, KV11, KS08b, KH09, Kul09, KKGT03c, KKGT03a, KKGT03b, KS10b, KZ10a, KR12, KJS01, KR06, KSG^{+}07, Kur07, Kur10, Kur09, Kur11, KCO04, KLT09, KMR12, Kut98, KŽ10b, KPW99, KCK08, KCG09, KL02, KS07c, KMKP11, KŚHM11, KDWW10, KFfKL11, KS10c, LLT07, LH11, LJS08]. International [Lac10, LV12, LRS03, LLS00, LO02, LCD^{+}09, LGK^{+}04c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, Lag05, LC06, LYL11, LZL11, Lai03, LOX99, LPT07, LK10, LM98a, LSV06, LVS08, LS10a, LLW06, LSS02, LLMT04, LdRV94, LSW00, LSW01, LSW02, LSW03, LO11, LN01a, LN04a, LPOA^{+}06, LNS06a, LNS06b, Lau01, LLCL05, LB05, Lau00, LS04a, Le 94, LZ08, LP97, LOR10, LLP99, LS95, LM98b, LT00, LBP00, Lee01, LMW01, LV02, LB03, LL03, LhK03, Lee04, LLWL04, LSgL^{+}06, LTW06, LL07, LKK^{+}07, LYY07, hLhKcFŚ09, LLPH09, LC09, LN09, LBM^{+}10, LYBU10, LCY10, LH10, LMMN10, LW11, LTL^{+}11a, LTL^{+}11b, LHŚ11, Lee11, LHKŚ12, gLPZ^{+}12a, gLPZ^{+}12b, LY12, Leh11, LWDM12, LOST04, LT05, LOTY05, LTÅ^{+}11, LM06b, LdQ07, LGG97, Leo04, Leo09, LIJ97, Leo00, LS88, LO98, LNE06, LVP04, Leu08, LM09, LS05, LM08a, LwCM00, LLLL08, LLC09a]. International [LLC09b, LPC^{+}11, Leu03, LW09, Lev98, LSY^{+}06, LPZ12, LSE02, LSHB07, LPH09, Li98, LWF04, LSDN04a, LSDN04b, LST^{+}05, LOS05, LBO05, LZhL06, LFIM07, LLIH07, LJJG07, LZS^{+}08, LSYO08, LKZ^{+}08, LHDD08, LFP^{+}09, LMA09, LFJI10a, LFJI10b, LJS^{+}10, LCB12, LPR12, LNW^{+}12, LEP^{+}10, LCF^{+}04, LSTZ12a, LSTZ12b, LN03a, LP10a, LFK^{+}02, LL04, Lin07, LTW11, Lin12, LMP04, LRL98, LN03b, LP10b, LPH00, LMCH12, LNS11, Lio04, LM03, LSK08, LKT08, LMWY04, LMW08, LMW10, LMW12, LR05a, LG95, LN01b, LCB97, LYhL^{+}01, Liu01, LCY03, LLL06, LH06, LL06, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LMA08, LLO08, LWYN09, LR09, LBDW09, LLWL09, LDXZ10a, LDXZ10b, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, LC11, LSIT11, LW12, LMC12, LLTY12, LK04, LS04b, Llo00, dVLKR^{+}08, LM04, LPR08, LMT10, LK07, LK09b, LRH10, LPA00, LPZ08, Lom97, Lom93, LN04b, LST^{+}12, LWRdR11, LPK05, Lop06, LH08, LT11b]. International [LHK02, Lor01a, Lor01b, Lou92, Lou94, LHJ08a, LHJ08b, LHJ08c, LS06b, LZ00, LZ05, LMM10, LZK11a, LZK11b, LZK11c, LRS98, LP08, LGS09, LR05b, LM08b, LCG97, LS12b, Luk97, LS09, LTT07, LV07b, Luo04, Luo06, Luo07, Luo08, Luo09, LSW^{+}10, Luo10, Luo11, Luo12, LMB08, LO88, LK00, LS10b, LHH11, LIH99, MJYT06, MRN07, MOP11, MOH^{+}12, MWD06, MRC12, MPK04, MT09a, MDY^{+}10, MDHB11, MS12a, MCK^{+}06, MKPV05, MMD05, MCKB06, MWYF10, MTT98, Mag03, MSB12, MP98, MMMP12, Mai00, MP06, MSG07, MG08, MBG12, MF11a, MF11b, MMV06, MT09c, MRK05, MNS05, MRvM07, Mal97a, MP03a, Mal02, MBS11, MJ04, MW91, Mal99, Mal01, Mal03, Mal07, Mal09, Mal11, MZ99, MBT^{+}03, ML04, MSP^{+}09, MZ07b, MSZ08, MNZ09, Man75, Man09, MS10a, MBW08, MY06, MMBND07]. International [MS03a, MSS11, MPdSM07, MT04a, Mar11a, Mar11b, MSS12, MS96a, MY01, MS06a, MS10b, MS10c, MS12b, MWY01, MVW09, MR01, Mar05a, MSZ90, MRŠ03, MMP03, MMV04, MPB05, MVC07, MSZ09, MVL11, MSY08, Mar11c, MdRIR08, MBH^{+}09, Mar01, MSS07, MVCM10, Mar93, MB01, Mar03, MNW01, MA03a, MF01, MVMP^{+}04, MOFM10, MLM90, MVOF08, MS05, MSW07, MR12a, MBVTY11, MDY07, MRZ09, MWZ10, MMDN12, MMP07, MTV09, MPT10, MKT98, MT07a, MMOS99, Mat01, MM03, MM07, MM09b, Mat02, MZ04, Mat09, MSWZ06, MTO99, MF08, MN02b, MMR91, MS03b, MPR07, Mau89, Mau96, ML11, MZCC09, MT97, MP09, MRY^{+}11, MP03b, May93, MST95, May01, MW97, McA00, MRPM10, MRP11, MPR12, McC97, MBP^{+}11, MG07a, McG08, MW09, MB06, MMM10]. International [MPvH04, MS96b, MFG12, MEB05, MNPW07, MO08, MTP^{+}02, MTS^{+}03, MT^{+}03, MT05a, MTH05, MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MTH08, MDH09a, MDH09b, MHD09, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, MZC06, MMO^{+}11a, MMO^{+}11b, Mei08, MT08c, MRSV07, MC94, MCA^{+}07, MH04, MSS03a, MS02b, Men07, MSW02, MLTC11, MAB^{+}02, MM10, MR08, MAFS08a, MAFS08b, MA11, MCSN10, MT02b, MJ07, MB09, MN12a, MN12b, MM04, MD96, MSH08, MS99b, Mid01, MMG^{+}01, MKZ02, MKZB04, MKZK06, MKZK08, MKZK10a, MKZK10b, MKPZ12a, MKPZ12b, MMMO96, MR07b, MMU96, Mil12, Mil00, MPPU11, MCSY07, MS12c, MMS93, Mio90, Mio93, MCP93, MS95, MMMD97, MdPA98, MP01a, MP01b, MA03b, MA03c, MÁ05a, MÁ05b, MÁ07a, MÁ07b]. International [MFÁ^{+}09a, MFÁ^{+}09b, MGH09, MV12a, MNS06, MS06b, MV12b, MKR^{+}04, Mit05, Mit06, MS08a, MKR07, MMK^{+}05, MS00b, MNLM06, MOT12, MMT08, MT99b, MK01, Moh91, Moh97, MKB^{+}09, MS00c, MPV01, MF02, MS08b, MTC^{+}08, MVA01, MAFSS04, MS96c, MvL10, Mon01, MMM05, MTU09, ML95, Moo01, MM97, MS11a, MF06, MG07b, MR12c, MDP04, MDBF01, MDPQA09, MDPQA12a, MDPQA12b, MMC05, MBS05b, Mor06, MK03b, MBJ^{+}07, Mos01, MMH07, MK12, MNS95, MSS04, MSY00, MS10e, MS02c, MR11, MSX10, MAR05b, Mue01, Mue06, MCdS^{+}08, Mug97, MOTN07, MTNL12, MC98, MPS04, MCER10, MSUB04, MSKY10, MSR99, MZ06a, MdSC09, MGSM12, MBC^{+}97, MV06, MA07c, MYS10, MMM11, MNW98, MNTY05, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d]. International [MS11b, MV07, Mur99, MBH^{+}11, MJL02, Myc95, MZ06b, NP02, Nad99, NDB94, NL01, NWL06, Nag90, Nag95, NDFMS97, NSM00, NW01, NNS03, NPPDG06, iNR08, NNPW12, NcC99, NR07, NYHO07, NZZ11, NMPM08, Nao07, NR99, NdMO03, NZS^{+}06, NSS^{+}07, NO11, ND09, NJ10, NP12b, NZ06, NHJ04a, NHJ04b, NHJ04c, NKPH08, NV08, NT92, NM94, Ng79, Ng93, Ngo09, NKN^{+}05, Ngu06, NGHJ07, NJHJ08, NBGA08, NKC09, NLŚ10a, NLŚ10b, NP10, NKJ11a, NKJ11b, NHJ12a, NHJ12b, NW97, Nic86, NCBT04, NT98, Nie99, NS00, NR00, NE02, NF07, NdG04, NWHR06, NV01a, NRWI09, NV01b, Nie03, Nie05, Nie94, NRR12, NBK04, NW10, NR04, Nik05, NR06, NCJK08, NQL06, Nip98, NY93, NK99, NPS06, NP11b, NLS08, Niv98, NT03, NABK12, NBC^{+}06, Nol81, Nor04, NSFB10]. International [NR08, NS99, NVSB^{+}07, NS90, NG10, NvRSS11, NBM09, NSL09, Nür03, ND04, Nyb98, NH03, Nyb08, NSS03, O’B12, O’H98, OROD07, OSR^{+}02, ONC^{+}11, O’S03b, OB07, ODP^{+}07, OO12, ONPR07, OT08, OFR09, OLL10, OS04, OiDP98, OKHH^{+}08, OKS02, Oka03, OYM07, Oka00, OW07, OA07, OP97, OZ01, OMMSP06, Oli07, OPS11, Ölv10, ORB^{+}09, OTZ01, OPT01, OPP04, OSV11, Omi00, Ong05, Ong10, Ong11, OdQ12, OSC^{+}08, OKdQ09, OA93, OYA04, OK96, OWPY03, OD12a, OKBMB10, ODV08, OR08, OF15, OV09, OSRS07, OST11, ÖRH12, OZ09, OZ11, Pac05, PAV^{+}10, PZ07a, PR08, Pag98, PHZ11, PRV11, Pai09, PV07, PPP07, PS02a, PKM04, PKKG06, PHJ03b, PHJ03a, PGM98, Pal00a, Pal03, PP01, Pal01, PR07, PT09, PJC^{+}12, PDH99, PDH01, PDHS03, PAD^{+}08a, PDML11, PPCS12, Pal00b]. International [PS09a, PCH^{+}06, PAD^{+}06, PZE^{+}08, PCN10a, PCN10b, PCN10c, PCK^{+}12, PS97a, PKM^{+}10, PDSD10, PSDS11a, PSDS11b, PDSN12, PP12a, PZK07, PZ08, POP09, PH04, Pap95, Par02, PA09, PR11, PBK^{+}10, PSST07, PV00, Par98, PCA^{+}09, PZYS12, Par09, PMÓ06, PMMR06, PSS^{+}10a, Păs09, PB07, PSPH^{+}10a, PSPH^{+}10b, Pat02, PdRM^{+}12, Pat03, Pat11, PMH09, PSE10, PFD^{+}10, PS12a, PH10, Pau03, PPJRN^{+}10, PT08a, PAD08b, PW92b, PMT09, PMT10, PZ96, PRK12, PMA^{+}07, PR05, Pel07, PV02, PW09, Pel11, PH97, PH11b, PA03a, PvES12, PP12b, PS96a, PRT06, PH02, PD04, PF06, PF08, PF10, PFM12, PpWR96, PP99, Per01, PBH03, PR03b, PI05, PS07a, Per07b, PS08a, Per09b, Per11b, Per12b, PT98, Per10b, PT92, PT08b, PCS^{+}07, Per11c, PS04b, PSD12, PK03, Pet03]. International [PW03, PU04, PVTG07, PSM10, PRH10a, PRH10b, PTZ03, PVP05, PY05b, Pet92, Pet02, Pez03, PHH10, PNB04, PB09, Pfe94, PS03, Pfe06, Pfe07, Pfi00, Pfi01, PTM^{+}11, Pic01, PMD90, PMD92, PMD94, PMDA96, PMD97, PMDK00, PMWO02, PSN95, POS00, Pie08, PZ09, PG01, PBK03, PK12, Pin95, PH06, PP03, PT99, PRdRW09, PB08b, PDG92, PS11a, PRJ95, PKSS02, Pla96, PK00, PHOK10, PLG06, PVV10, PP11, PMC06, PJ12, Pol86, PS98b, PS09b, PdA^{+}11, Pol97, Pol99b, PZH96, PC00, PBRB08, PBS08, PLK^{+}12, Por98, PR00, PS07b, PG09, PISW05, PT91a, PT91b, PZ07b, PLI08, Pre00, PT06a, Pre09, PT11, PF07a, PWY08, PLMV12, Pri03a, Pri03b, PQ05, Pri06, Pri91, PPM09, PM09, PRR05, PCA02, PPH07, PR98b, PRR94]. International [PR97, Pro96, PBW05, Pro11, PS12c, PS11b, PW07, PW08, Pue07, PSS98, PPF^{+}00, PP04, PS09c, Qin10, QQ11, QOZ01, QGZ03, QIW07, QMW09, QSWL11, QLXL07, QSBC98, QBY^{+}12, QS00, RKS89, RCC12, Rah04, RF10, RiN12, RPR08, RM09, RWSS12, RWA06, RSV07, RMDS10, Raj09, RFBW12, RW11, Ram01, RR08, RLL04, RS10, RCA90, RS06, RR12, RUL04, RRU05, RA12, RM02, Ran93, RD01, RFZ03, RVY05, RS00, RS05, RCJ^{+}08, RT12, RZ91, RK93, RZ94, RS99b, RO00a, RTZ08, Ras01, RT07, Rau11, RMFG06, RRBE09, Rau04a, Rau04b, RKM^{+}12, RR98a, RWY97, RV10, RDPMA^{+}11a, RDPMA^{+}11b, RR01, RBG09, RD12a, RA00, RTB02, RST09, RK06, Rei89, RBF06, Ret97, REF^{+}06, Reu99, Reu01, RMS^{+}05, RSS06, RL06, RN11, RM04a, RMB10, RS12, RGW01, Ric98, RYG11]. International [RMTW11, RABR03, RF99a, RL11a, RPBP06, RSL01, Rob11, Rob06, RR04b, Roc07, RL11b, RH01, RCLE05, RK02, RKW04, Rol97, Rol98, Rol99, RV02, RBC93, RV11, RCS10, RHD08, RB10, RJS^{+}08, REPH06, RS03, RT10, RF12, Ros03, RT09, RBSL06, RPZ97, RH98, RFBD09, RS01, RV03, RSW12a, RO00b, RM04b, Roy05, RS09c, RD12b, RJL^{+}08, RDK09, RG11, RB00, RGLA09, RBB^{+}06, Rus00, RZ12, RSTZ04, RTZŻ06, RTZZ08, RST^{+}10b, RST^{+}10a, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, RKNV09, RME04, RS09d, RSW12b, SST00, Sab10, SBV^{+}11, SS07a, SPBP08, Sad13, SP09, SNČ^{+}12a, SS08a, SNY06, SN08, SO07, SW04, ŞS05a, ŞSW07, SKS02, SCH^{+}09a, SKS^{+}06, SM11a, SS11a, SSVW12, Sak91, Sak01, SFS11, SGSV10, SL11, SdSMO12, SGPP06, SL12, SS11b, SYNC11, SÇS12, SS09a, SS11c]. International [SW09a, SYMA09, STP^{+}10, SZR^{+}08, SPCG07, Sd98, SVS02, SKR11, Sas05, SHM^{+}10, SNE08, SMBQ07, SPH07, SCKR10a, SCKR10b, SMB03, SHØ06, Sch10a, SSRA98, Sch04a, STT11, ST00, ST08, ST11, SS01, SKP07, Sch93a, SH12b, Sch93b, Sch03a, SUW02, SKP10, Sch11, STL11, SCM88, Sch06, SCHS07, Sch09b, Sch12, Sch74, SR07, SEG09, SF11, SE94, SB07, Sch01a, SMH^{+}12, SDR^{+}00, Sch10b, SJW^{+}12, SvS00, SC11, SS05b, SNRG11, ST12a, SW03b, SWAF95, Sch07, SSN08, SNZ08, SS09b, SVV12, SS06a, Sco02, SLZH07, SVS08, SZ92, Seh97, Sei07, STT04, SHGB05, SHW09, SP08, SZC^{+}03, SG10, Sen10, SO09, SSJR10, Ser12, SC09a, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, SG09, SEO^{+}09, SPK01, SMP07, SHX^{+}06, SW09b, SW08, Sha86, SCL^{+}06, SLL^{+}07, SYY^{+}08]. International [SPÖ^{+}10, SBC^{+}94, She01, She03, SSD^{+}08, SS06b, SJ06, SvADS07a, SvADS07b, SvADS07c, SvADS07d, SKK08, SKPG12, SKSA12, SILS05, Sho84, Sho05, Shu07, Shu09, Shu11a, Shu11b, Shv06, SF08b, STA^{+}11, SCG98, SBD03, SCR06, SA06, SPON08, SAG10a, SAG10b, Sie86, Sif90, Sil01, SRW98, SMK01, SSAP05a, SSAP05b, Sio10, SMPU98, SEL09, SB12a, SMEGA10, Sir12, SL99a, Ska01, SYP^{+}05, SA12, SLL01, SBH98, Slo99, STDH02c, STDH02a, STDH02b, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, SCH04c, Sma08, Smi01, Smi05, SS07c, SS07b, SS09c, SS11d, Smi12, Smo97, SLSW12, SPO11, SKP00, SKP02, SKP04, SKP06, SHKP08, SHKP10, SHKP12, ST07, SSM08, SCZ10, SPSV^{+}04, SL99b, SOP12, SAUL11, SK97, SDP97, SZ00, SK08, SBM11, Son00, SLLS03, SEN07, SEN08, SMF^{+}11, SMMG01, SQL10, SPB00, SMK02, SLPM96]. International [SLKL09, SH07, SS08b, SMK06, SR11, SH09, Squ09, SF08c, SRY07, SB12b, SS06c, SSU12, SS99, SKCK07, SHL08, SvdMOP06, SDH03, ST12c, Ste98, SL03, Ste95, Ste01, SW97, SDK^{+}08, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, STWS07, STTP02, Ste99, SWB03, SS09d, Sti90, SG07, SBF08, SSKR02, SCKR03, SRT06, SHS10, SD10, Sti94, ST01, SP12, STY^{+}07, SWMM06, SFV^{+}11, SGD09, SKB98, Str99, SS10, Str96, Stu02, Stu08, Stu90, ŠPTM00, SDM05, SZ11, SBH07, SBH09, Stü09, SRL12a, SRL12b, SRL12c, Sub05, SV01, SHRM06, SZT^{+}08a, SZT^{+}08b, SNSR06, Sur02, SLM07, SHYEF08, SHUH08, SWSY11, SM09, SC00, SHO99, SBH^{+}04, Syv02, SV12, SKN05, SHŚ^{+}07, SKR^{+}10, SH11, TDB^{+}06, TR12]. International [Tah00, Tah01, TMLL09, TOOO07, TM09, TWQ^{+}12, TDWG11, TBE08, TT95, TMG10, TMdB11, TMS12, TSG00, TLF01, TST10a, TST10b, TSCW11a, TSCW11b, TSJ12a, TSJ12b, TFK11, TWYhL01, TLZ^{+}08, TKCW11a, TKCW11b, THL11, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, Tan11, Tar91, Tau08, TM99, TC99, TN03, TBKO10, TY11, TdLdOQ08, TV94, TYZ11, TTM10, TDS94, TD07, TFE^{+}09, TS10, Tha96, TBC07, TNA^{+}10, TKSN11, TCYZ12, TPCB12, Thi10, TW12, Tho99, TG00, TGH06, THX^{+}07, TBH03, TSS11, TS87, Tis02, TBJ02, TN09, Tiu00, TT06a, TQYX11, TLJ06, TBF^{+}09, TO08, Tok07, TP09, TC98, Tom92, TNP^{+}07, TG02, TT06b, TN04, TNM05, TNVDF06, TNY07, TN08, TNI09, TND10, TNYL11, TSK92, TCJ10, Tou94, Tou96, TTF07, TG10, TL02, Tra04a, Tra04b, TGAH11, TKHS05, Tre09, TST09]. International [TZS00, Tri98, TMN09, Tro11, THR03, TCLH10, Tsu08, TSKGB04, Tuc01, TtHVS06, TWSN07, THT03, UPB07, UWB^{+}06, UMN^{+}08, UBM02, UBW06, Uus06, UDF06, VvS99, Vah09, VJKT00, VKV00, Val94, VGP08, Val06, VJ11, VY93, Van97a, Van02, VW07, VAR12, VM99, VMvN12, VW05, VV03, VWH06, VBH07, VEG^{+}11, Vau98, VY01, Vau05, Vau06, Vau08, VFC08, Vck99, VJAA03, VA07, VRHJ09a, VRHJ09b, VA95, VG12, VDPD96, VC03, VMBMN04, Vid12, V^{+}06, VKJ03, VZM^{+}10, VKMT11, VDÉ^{+}12b, VDÉ^{+}12a, VSMS12, VBS^{+}10, VMB^{+}12, VW08, VV07, VD12, VT97, Vit10, VZ99, VGB11, VS05, Vog01, Vog88, VSTN06, Vol93, Vor02, Vor08, Vos03, VLY09, VAM06, VASP09, VKG12, VWY97, VWY01, Vyt91, WF98, WK08, WAS06, WT96, WR01, WRP07, Wag08, WSS04]. International [Wal01, WKP03, Wal04b, Wan98, WYL01, Wan01, Wan03, Wan04, WLY05a, WLY05b, WLY05c, WJ05, Wan05a, WYZ06a, WYZ^{+}06b, WJS^{+}06, WPSY06, WLC^{+}06, WPRW06, WmCL07, WPR07, WPR08, WLGB^{+}08, WFZL09, WGMP09, WDGL10a, WDGL10b, WYSS10, WGLL10, WZD11, WLO^{+}11, WS12, WYP12a, WYP12b, WJL^{+}12, WZJX12, WZH^{+}12, WB00, WT07, WZ03, WZH^{+}07, Waś96, WY99, WZ09, WW03, Wat00, WK02, WR07, Wel01, WC98, WMR00, WLP^{+}09, WMS04, WL04, WHG07, WVD02, Wes01, WPM06, WPM07, WBD08, WMdLA10, WG11, Whi98, WCK11, WNE99, WTM^{+}02, WB09, WH95, Wie99, WP10, Wii04, Wii05a, Wii05b, WBW94, Wil01a, WR09, Wil99, WM03, Wim03, Wim04, WSGA06, WSG07, WSF08, WSJT09, WCJS10, WJP07, WFB12, WR03, WL03, Win04, WNL05, Win11, Win09, WDKK07, WZA10, WN96, WPH03, WKB04, WHR10, WLS89, WHS01, WPD04]. International [WPZA11, WZ11, Wol95, Won01, WK06, WMB10b, WMB10a, WY98, WL93, WCBD08, WWC02, WANL08, Wri03, WVDS05, WCGB05, WYX08, WLRL08, WD10a, WD10b, WKD08, WDPW02, WDPW04, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, XW11, XCHZ11b, XCHZ11a, XSA^{+}12a, XSA^{+}12b, XLKZ12, X^{+}12a, X^{+}12b, XYM^{+}07, XKGS12, XXX^{+}09, XBS^{+}10, XCR11, XHXL08, XLHX08, XL11, XZCL07, XYZU11, XBM12, YBM87, Yah11, YNB08, Yak00, Yak02, Yak04, YN06, YD98, Yal10, Yam08, YGGJ04, YEY04, YJ04, YJMU06, YJS^{+}06, YW06, YDW08, YCC^{+}08, YHR^{+}09a, YHR^{+}09b, YYI^{+}09, YMHH10, Yan11a, YGE12, YLW^{+}07, YSP^{+}10, YRWS11, YYS^{+}12, YLS^{+}12, Yas00, YŞ03, YI12, YLWY06, YKF^{+}06, Yi06, YAF^{+}02, YWG04a, YWG04b, YTC^{+}07, YWRS11, YCB12, YJKK06, YJKSW10, YGGÖ05, YM01, Yor08, YC11, YSL12, YHV12, YMY^{+}10, YSR^{+}06, YY09, YKM06, YKMJ09, YAKMJ10, YAKMJ11, YAKM^{+}12]. International [YP01, YKL06, YHZ09a, YHZ09b, YHZ09c, YGL^{+}10, YLC^{+}10, YKU11a, YKU11b, YYH^{+}12, YY01, YZCW07, Yun02, YDKM06, YLL09, ZBW94, ZSSM03, ZRP^{+}09, ZRP^{+}11a, ZRP^{+}11b, ZK06, Zam08, Zan00, ZTTY12, ZBRL12, ZCCL10, ŽL08, ZCRL08, ZPT^{+}06, ZS00, ZJ04a, ZZ04a, ZGY04, ZS07, ZOCJ07, Zha07, ZSM^{+}09, ZPTI09, ZLK10a, ZLK10b, ZZP^{+}10, ZCDT10, ZS10, ZO10, ZHLW12, ZG03, Zhe02, ZJL06, ZSO99, ZRTS03, ZLY^{+}07a, ZLLC09, ZCGH11, ZYH03, ZJXC03, ZNCF03, ZSP^{+}04, ZYB06, ZZTG07, ZYDL09, ZY10, ZZLL10, ZF05, ZY01, ZBH12, Zim92, ZJS^{+}02, ZT04, ZVU99, Zit01, ZBY07, ZBG98, Zuc03, Zui99, dCdF^{+}08, dA09, dBBGdR03, dBBG04, dBBGdR06, dBBGdR07, dBBGdR08, dBBM09, dBBHL10, dBBC02, dCVK07, dFEN04, dGEK11, dGN12, dKIM^{+}07, dLZ12, dMB05, dMS06, dMLG03, dMS09, dMGFS11, dRLP98, dRWK^{+}10, dSADM07a, dSADM07b, dSDL04, de 98, dS02, dSOSR06]. International [dS11a, dVGR04, dPMA98, dlBP08, tCZ07, tTVH^{+}12, ter97, tHBP08, vBC08, vB97, vDLJ11, vEGW09, vES10, vEGSW11, vEDA04, vHV08, vHH09, vKS12, vL88, vL89, vLS91, vL94b, vLWH^{+}00, vLS11, vO04, vSN01, vSH06, vdPW10, vdBHOS09, vdBK05, vdBGG10, vdBV01, vdHI99, vdHHsHD06, vdHCD07, vdHXMW08, vdHS10, vdHIP11, vdHP12, vdAB03, vdAHW03, vdL98, vdL01, vdL02, vdL04, vdMBD08, vdMvdT08, vdMS11, vdPS08, vGH96, vzGIK08, AS94, ADG^{+}08a, AMW07, AH94, A^{+}08, ABLM05, AN95, AMSM^{+}09a, AM91c, ANS^{+}09a, ANS^{+}09b, AH96, AS06, ABD04, ADW05, ADL05, Ang97, AM04, ACJT07, Asa96, ACL^{+}04, AGL06, ADCR89, ADL09, Aya95, AB97, BV05a, BCG02, BPSP06, BLPW03, BL05a, BCKK05, BMSMS04, BM05b, BBH^{+}05, BUG93, BS76, BB79]. international [BMH05, BBKP06, Bec05, BHGP02, BIRT05, BCT06, BB05a, BR06, BC04a, BB05b, BGD^{+}05, Bha95, BJ86a, BMW93, Bir07, BBP93, BTPPS05, BGU94, BGU95, BC05, BMM07, BC07, BP04b, BCKT05, BL06c, BJS93, BW05b, Bro92, Bro94, BV05b, BRR97, BvADS08c, BC04b, BS06c, BR96b, BT06, BTF05, BS06d, BFKO05, yCW96, yCCZ07, CIM^{+}05, Cal86, CCP96, CDP^{+}05, CC05, CYGL04, CNX05, CP06a, CO94, CR05a, CTT93, CPGL05, CSV05, CD05, CAZ05, CG88, CW89, Coh96, CM05a, CFR92, CN05, CP05, Cou93a, CKL05, CM05b, DWSK05, Dam99b, DG96, DV05, DN04, DR05, DDFM05, DS05b, DE05, DIMV11, DGMS97, DLOS05, Del97a, Del97b, DL04a, DD06, DHX09, DJ90, Des94, DWML05, DR81, DPA05, DF96, Die02, Dil94, DJS91, DH06, DRS05, DOSY05, DBB^{+}04, DvV96, DLW06, Dov05]. international [DL95b, DG05b, Dyc94, DHSH96, EH95, EFM04, Ehr04, ELM05, EL04, EHSH92, Esi93, EW05, Eta05, ER05, FS05, FKP^{+}05, FLS06, FRS05, FFRS05, FHRR05, FMO05, FPMW06, FT04a, Fra05b, FP05, FK95a, FKKB^{+}05, Fri04, FW92, FLMS05, FLS05, FG95, GO94, GG05a, GP05, GHM05, GG05b, GMV05, Gan96, GL97, Gel05, GWS05, GM04a, Gia04, GL05, GS05, GPPW08, GLW88, GK09, Gra05, GHMT05, GT05, Gru97, GAN05, GM05, Gut94, HMRT05, HZ05, HOQ97, HH04, HS05a, HR96, HHM97, H^{+}05, HK94, HN06, HCS^{+}05, HSSY05, HT05, HC05, HPR05, HIS05, HS97, HML97, HS98b, HW06b, HKNW07, Hsi95, HM04, HW05, HM05b, JPBT05, JP05a, Jan93, JSZ95, JKB04, JS05, Jen92, JNS05b, JJ03, JL97, JRJ05, Joh97, JP94, Jou94, JK05a]. international [JK05b, KSHS01, KG74, Kal97, KC05, KJ88, Kar94, KKM04, KN91, KIRK05, KD05, KHJ05d, Kim05a, KSW05, KW90, KL92, Kir96, KKKN05, Kit88, Kob96, Koh06a, KS05b, Kot86, KS04c, Kra05, Kui92, KK95, KA05, KBC05, LY04, LM93, LS92, LSW98, Lau93a, LMS96, LD97, Lei95, LBCO04, LLC^{+}05, LA94, LM97, LWD05, LVW05, LV05, LCHS96, LSS^{+}04, LY05, LMV95, LKC93, LMW06, LS89, LR94, LSL04, LA05, LQO04, LD05b, LD05a, LG05, Luo05, LKZCS05, Mac06, Mai90, Mal95, Mal97b, Mal05, MNT95, MR97a, MST92, MD04, MBS05a, MS00a, MMT97, MMP05, MPPJ^{+}05, MSW05, MH05, MXZ05, Mel86, MM82, MM96, Mic90, M^{+}06, MR97b, Mol95, MAA91, MR95, Mor89a, MT05b, MSS05a, MBA05, MSH05, MB87, MC05, Nak92, Neh88, NOOS05, Nic75, eOP05, OGM05, Oli00]. international [OFK05, OSvL01, Ott87, OPKR05, PBB05, PTRH05, PS04a, PVV05, PCG^{+}05a, PGA^{+}05, PS98a, PFCS05, PC05, PF05, Pat90, PY05a, Paz97, PT05, PSO05, PV04, Pre95, QMLW05, RMR05, RM96, RS97b, RMS04, RR03, Rei95, Reu97, RT95, RB91, Rob80, RV05, RBS93, RSvdP05, SV97, SH91, Sch92, SP90, Sel95, SDH05, SA96, SLB05, SLL^{+}06, SJT05, SBG04, Sma05, SKY06, SC96, Sta95, SS06d, SZ05, Ste94, SLP05, STLK05, SvASD05a, SvASD05b, SvASD05c, SV05, SA04, SC05, TDBP01, TPB79, TLC06, TTP05, Tis94, TT05, TEB09, TMP04, TYNM05, Urz05, VZ06, Van05, Van89, VU05, VH04, Vor92a, Vor93, WS02, Wal04a, WCO05a, WCO05b, WCO05c, Wan05b, W^{+}12, WPM05, WTGA05, WGD05, YZZ^{+}05, YAL^{+}05, YRMD05, YBL^{+}04, ZHF04, ZLJ05, ZJ06, ZGT05, ZOM05, ZLDB05, Zip92, ZS05b]. international [dBBGdR05, vB05, vEHD05, vWN99, vdABCC05, vP93, von96, AFKL04, AS07a, BU01, DCM82, EN74, MSA99a, MSA99b, PR84, Rob80, Ytr06]. Internationale [Sch74]. Internationalization [Ayk07a, Ayk07b, Ayk09, Rau11]. Internet [ALS^{+}11, ASE^{+}07, AH02, ÁCD^{+}12, Ber02a, BGPTF10, BMM07, BRU06a, BHU03, BCKT05, BCF06, BBB06, BKB03, CA08, aCF06, CW11, Cha05a, Cha02a, CCNY10, CJ05, CJ06, Chu03, CMS11, CRS00, DYCD09, DPLW08, DY10, DFT09, DGG^{+}11b, FS07b, FLF^{+}08, GVCA07, GM04a, HHLS08, HTMC09, JM07, JM10, JM02, JMK09, Kah03, KS03b, Kim01b, KW98, KMMT12, LLMT04, MCK^{+}06, MPP08, MBVTY11, MTP^{+}02, MTS^{+}03, MT^{+}03, MT04c, MT04d, MTC04, MT05a, MTH05, MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MTH08, MDH09a, MDH09b, MHD09, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, NO11, OS09, OKBMB10, Pal01, PPPM10, PA09, RR12, RST09, SILS05, Smi01, Smi03, SPSV^{+}04, SSKR02, SCKR03, SRT06, UBM02, VS09]. Internet [VT97, VJ99, WL04, ZSP10, vdBHOS09, KS05b, ASP01, ABK12, BCB99, BU01, CEK11, Chu03, DY06b, DG07b, HL00, Kou11, LT01, Leo09, PZ08, Sab10, SMK06, Sto04, X^{+}12b]. Internet-Based [Ber02a, CMS11, DYCD09, WL04]. Internet-related [Kim01b, SILS05, Chu03].

INTEROP [MTH05]. INTEROP’99

[Vck99]. Interoperable

[BGH06a, EM06, EK10, FR11, GH12b, IR07, MT08c, SO09, SDH03, BBH^{+}05, KA05]. Interoperating [Vck99]. Interplay [FSdlPT11a, FSdlPT11b, Gol11b, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b]. Interpretation [BH10, CP07, Cou05, EN06, JS11a, JMO09, Kon04, KR12, LPZ08, SL03, Cor02, Zuc03, Beh03, Mul93]. Interprocedural [Kno98]. interval [CKPB07, Nic75, Nic86]. Intervention [AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, BHH04b, BHH04c, DK02a, DK02b, DG05a, EP03a, EP03b, FMP11a, FMP11b, FMP11c, JNPV10a, JNPV10b, JNPV10c, LNS06a, LNS06b, MDY^{+}10, MAFS08a, MAFS08b, YHR^{+}09a, YHR^{+}09b, DDJ00, DG05b, TC99, WC98, NV01a]. intervention-MICCAI [NV01a]. Interventions [AJNJ12, LMCH12, MDHB11, NJ10, TY11]. INTERWORKING [RS00]. INTETAIN [MSW05]. Intra [SPO11]. Introduction [Ber12, BM03b, DT07, Neu04, BWT95, CJE82, Woo80]. introductory [GH92]. Intrusion [DMW00, JSK10, JVA04, KJB09, KLC07, LKT08, SBM11, VKJ03, WVD02, ZK06, LMW01, VZ06, JK05a]. Intrusions [BSC12, BL06d, FB09, HS07, HB11, KJ10, Zam08]. INTRUST [CY10, CY11a, CYZ12]. Intuitive [GT05]. invariance [MZF94]. invariant [Rei93]. inversion [Yel88]. Investigation [Pup10]. Invited [AT07, APN08, ABFR07, BE06, BSvRW08, BSvRW09, BBvRL10, BFGM05, BHM^{+}04, BHMS07, BB06, BDDS05, BDDS06, BDDS07, BR07b, BHJY07, Buc85, CCS04, CFP^{+}09, CDM04, DDE04, DSRW08, DDHP05, DPC^{+}07, Dig04, DGR06, DKN12, DFN10, DS07, EL07, EKP^{+}07, EFZKM07, EV09, EHMM09, FBKS03, FBSMS08, FPFR05, FST09, FPRS06, GOZ05, GGHS07, HPS09, HCF06, HPRS06, HNPP07, HKS09, IST07, IGN05, IFV07, JDMB08, Kap08, KCO04, LOST04, LT05, LMP04, LABM^{+}12, MPR07, MRPM10, MRP11, MK03b, MZ06a, Nie09, NLS08, ND04, OROD07, OPP04, OSV11, PZ07a, PMMR06, PPJRN^{+}10, PW09, PNB04, RPR08, RM09, RS99a, RtTMP10, RME04, SS08a, SSVW12, SNZ08, SA06, Squ09, TGH06, TT06b, VASP09, WPM06, WPM07, WBD08, WMdLA10, ZLY^{+}07a, ZYW08, dCdF^{+}08, vEDA04, vdMS11, AJS05, CSV05, DL04b]. invited [DRS05, Fre98, GOW03, GS05, HPR05, KS05a, Li96, MPPJ^{+}05, MSS05a, MBA05, PDH99, PDH01, PDHS03, RMR05, SVV12, Van05, VT97, WWC02, vEHD05]. IOT [FLF^{+}08]. IOTA [NY83]. Iowa [BMP90, HS98b, Rus00]. IP [ABLM05, MMD05, PMÓ06, SSN08, APS08, MB01, Mar03, MNPW07, NSL09]. IP-based [MMD05]. IPAW [FKM08, GF12b, MMM10, MF06]. IPCAI [AJNJ12, NJ10, TY11]. IPCAT [LST^{+}12]. IPCO [ES10, FW07, GW11b, JK05b, LPR08, AG01, BN04, BBRM98, CS02b, CBW99, CMQ96]. IPDPS [FR00, Rol00]. IPEC [MR12a, RS10, TW12]. IPM [AS10, AS12]. IPMI [CS05b, KL07b, PPM09, SH11, BG93a, CH91, IL01, TN03]. IPMI’97 [DG97]. IPMI’99 [KSTP99]. IPMU [HKH10, BMYZ91, BMYZ95, VY93]. IPOM [APS08, MMD05, MNPW07, NSL09, PMÓ06]. IPPS [FR95a, FR96, FR97, FR98, Rol98, Rol99, RF99b]. IPPS/SPDP [Rol98, Rol99, FR98]. IPPS/SPDP’99 [RF99b]. IPSEN [Nag96]. IPSN [ZG03]. IPTComm [SSN08]. IPTPS [CR05a, VS05, DKR02, KS03a]. IPW [MT^{+}03]. IR [ACR07, BYdVZ^{+}12, BBMSD09, CFG^{+}11, GHK^{+}10, LMR^{+}06, LFL05, MOP^{+}08, CGvR02, MT04b, Seb03]. Iran [AS07a, AS10, AS12, ST02]. Ireland [AST05, BVO10, BR04, BS06a, BH06c, BW10b, BMJ88, BH95, BW01, BBK^{+}11, BLS02, BS11f, BGH06b, But10, Cag02, CFM09, CW01, CFG^{+}11, CF10, DB07, EKO^{+}04, ECV^{+}10, FSN06, Fos02, GGK06a, GMBM05, HKS06, HN06, HJMP06, KMF06, LKD08, MA03a, MH05, MSS05b, OROD07, OSR^{+}02, O’S03b, PP01, PMÓ06, Pol01, PdA^{+}11, RS09b, Ros03, RG11, SC98, SvdMOP06, Ver00a, Ver00b, WAS06, WR07, tTVH^{+}12]. IRFC [CHR10, HRdV11, SL12]. Irish [CF10, OSR^{+}02]. Irkutsk [Vor92b]. IROBOT [NSM07]. IRREGULAR [Fer96, Bil97, FR95b]. IRREGULAR’98 [Fer98]. irregularly [Bil97, FR95b, Fer96, Fer98]. Irsee [ANR99, ADMH04, ADM^{+}06, ADM^{+}08, Mol95, RABR03, SH11]. IRSG [CGvR02]. Irvine [GK02, JT09, XL11]. IS&N [MBC^{+}97]. IS-EUD [CDFP11, PRdRW09]. ISA [hKA10, PCA^{+}09, HL91]. ISAAC [Asa06, AiNOW11, CCP10a, CCP10b, DD06, DDI09, FT04a, HNF08, Tok07, Asa96, BM02b, DZ94, ET01a, Iba92, IKO03, LT00, LIJ97, Ng93, Sta95]. ISAAC’98 [CI98]. ISAAC’99 [AR99]. Isabelle [NPW02, Pau94]. Isabelle/HOL [NPW02]. isAI [NMM10, OBS12, OBM11]. ISARCS [Gie10]. ISAS [MRK05, MNS05, MRvM07, NMPM08, PRT06]. ISBMDA [MCKB06, OMMSP06]. ISBMS [BE08, BC10a, HS06]. ISBRA [BMSW12, BGPR10, CWZ11b, MZ07b, MSZ08, MNZ09]. ISC [BTMI11, De 07a, GLMP07, GF12a, KLB^{+}06, LZL11, SYMA09, WLRL08, ZLDB05, BM03c, CG02, DF01b, ZZ04a]. ISCA [VMvN12]. Ischia [AR10, BG10b, DGT10a, DGT10b, Fes10, GVT^{+}11, SMP07]. IScIDE [ZZZL12]. ISCIS [LSY^{+}06, YGGÖ05, ADK04, YŞ03]. ISCO [MMMP12]. ISCOPE [COT98, Ish97, MTO99]. ISCSLP [HMCL06]. ISD’99 [AS99]. ISDE [MHD09, MDH10c, MDH11a]. ISDF [M^{+}06]. Ise [GH08]. ISEC [KYT01]. ISFEM [DW03b]. Ishikawa [DMO98]. ISHPC [LJS08, VJKT00, VJAA03, ZJS^{+}02]. ISHPC’97 [Pol97]. ISHPC’99 [Pol99b]. ISI [KMR^{+}05, MZC06, YCC^{+}08, Che03a, CMZL04]. ISICA [CLKL09, CHKL10, KLZ07a, KCYL08].

ISIPS [GKL09]. Island [APS08, ACP05, AS10, Bai05, BC08, Bor94, BL10b, BBK^{+}11, BGR08, CY03, CCP10a, Cho02a, Cho02b, CSY09, CY11b, CM04a, DWSK05, DKC10, DKK08, FSM03, Fes10, FC93, Fuc10, GCK06, Had03, HK05c, JY12, Kah03, KPS12, KGL04, KLB^{+}06, Kim01a, Kim05b, Kim05a, KYL07, hKhLKŚ10, hKAS^{+}11, KR02b, Lee04, LLWL04, LYMM05, LYY07, hLhKcFŚ09, LY12, LY05, LD05b, LD05a, MPF95, MN12a, MN12b, Nik05, NCJK08, ONPR07, PV00, PAD08b, RG03, RF12, SW03b, SJ06, SRW98, SB12a, SKY06, SKCK07, STLK05, SHŚ^{+}07, TNI09, vdMBD08]. Islands [BHHK08, FSdlPT11a, FSdlPT11b, JJF03, MMMD97, MÁ05a, MÁ05b, Sio10, SCD^{+}10]. Isle [GL03]. Islet [Dan12]. ISMB [BS10]. ISMB/ECCB [BS10]. ISMDA [BH00, CGS02, CMM01, PBH03]. ISMIS [AMRŚ08, CFLR12, ERMS06, HMRT05, KRSR11, RRBE09, HRZK02, RZ91, RK93, RZ94, RM96, RS97b, RO00a]. ISMIS’99 [RS99b]. ISMM [SPO11, WR09]. ISMS [CM04b]. ISNMDA [BMSMS04]. ISNN [LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, SZT^{+}08a, SZT^{+}08b, WLY05a, WLY05b, WLY05c, WYZ06a, WYZ^{+}06b, WYP12a, WYP12b, YHZ09a, YHZ09b, YHZ09c, ZLK10a, ZLK10b, YWG04a, YWG04b]. ISO [TDB^{+}06, TDBP01]. ISO/IEC [TDB^{+}06, TDBP01]. ISO605 [HS12a]. ISoLA [MSS12, MS06a, MS10b, MS10c, MS12b]. Isolation [Lib07]. Isomorphism [Thi00b, Hof82]. ISOOMS [BU94]. ISOTAS [FM96]. ISPA [CYGL04, CPGL05, GYD^{+}06, M^{+}06, PCG^{+}05a, STY^{+}07, THX^{+}07, GY03]. ISPEC [BLW09, BW11, CDLZ06, CMS08a, DW07a, DBPZ05, KDWW10, RSW12b]. ISPRS [SRM^{+}11]. Israel [AS94, AVP07, AL01, AHW07, BHKR11, BB99, BD97a, Bih97, BZU07, BtHRBB12, CH09, DSM94, DL95a, DGR92, DSS^{+}07, ELS12, ES00b, EKM06, EK81b, ER12, FKK09, Gel08, GSLM12, Gru97, HG02, HBN06, KS08a, NZZ11, PT99, SZ92, TCYZ12, UBW06, Yor08]. Israeli [Coh94]. ISS [OBM11]. ISSAC [Gia89]. ISSADS [RRU05, RUL04]. ISSEP [HKV10, KM11, Mit05, Mit06, MS08a]. ISSS [FMY04, Oka03]. ISSSE [DF09]. Issue [GTM09a, GTWC09, GTA10, GTM10a, GTM10b, GTSS11, GTM11, GTP12, HKW11a, HKW^{+}11b, HKW^{+}12b, JvdA09, KOFJ09, NJA^{+}11, PBG^{+}10, PPdV12, SK12]. Issued [CKOS09]. Issues [ABF^{+}07, ADM^{+}12, AL11b, BRS^{+}07, BJSW09, CGG^{+}07, CM00, DV09, DO89b, DGDM^{+}11, EHMM09, EVV^{+}11, EEM^{+}11, Fed01, GLZ^{+}03, GCL^{+}06, HHL07b, HWLJ08b, KSFN08, KS07a, LSgL^{+}06, LY01, LGR^{+}04, OS07, SNY06, TC95, Yal10, AYB95, AY96, Can89, CRW97, Che99b, CGLR05, Chu03, DC01, GM00, Kim01b, LG94, RBS93, SILS05, Ste12, ŠPTM00, TLC00, VJ99, WGL96, DG00]. IST-AWSN [THX^{+}07]. IST/FET [PQ05, Pri03b]. Istanbul [BdBK03, CM10, DCC^{+}10, DBD^{+}10a, DFS08, EAES^{+}10, FHY08, GJTS06, HH10, JOD98, JP12a, Kay03, KMS12, LSY^{+}06, PRG10, SMD04, SGSV10, Sma08, ÜÇA10, Yal10, YGGÖ05]. ISTCS [DGR92]. ISVC [BBKP06, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BBP^{+}12a, BBP^{+}12b]. ISW [DMO98, POS00]. ISW’99 [MZ99]. ISWC [ACN^{+}07, AWA^{+}11a, AWA^{+}11b, BKH^{+}09, CDA^{+}06, GMBM05, PSPH^{+}10a, PSPH^{+}10b, SSD^{+}08, dCdF^{+}08, FSM03, HH02, MPvH04]. IT-security [FH01]. Italia [DGT10a]. Italian [AMT03, AMNT06, AHJ07a, AHJ07b, AHJ07c, AGS91, BM05a, BP07, BL03, BBD97, BGP00, BCP94, CFI06, CT03, CLP05, DP10b, Esp01, GS95, LM00, Len97, MT02a, PPS03, PS11a, RRR01, SC09a, Tor93]. Italy [AJNJ12, AZ96, ACP01, AT02a, AFP^{+}10, ABLM05, ABBFM11, ABBC10, ABDW10, AE05, AR10, ABD04, ACR07, AC11, ABM02, APW06, AH98, AMT03, AMNT06, AHJ07a, AHJ07b, AHJ07c, Apo93, AM04, AGI^{+}06, AGM03, AGS91, Arm02, ABD03, AL06, ACR12, AGL06, AM05, AB78, Aus86, ADCR89, Avg12, BM96, BTT90, BM05a, BMUV10, BSW95a, BGLT05, BFH07, BCKK05, BBM^{+}08, BG09, BP07, BW03b, Bat01, BCL03, BH08, BDG12a, BBB07b, BDCM04, BCST99, BC96a, BCD04, BDd09, BFBS11, BdR94, BCG^{+}11, BI03, BC04a, BB05b, BH07, BDZ08, BPZ09, BI11, BCP12, BU94, Ber96, BKW08, BM10, BVAT12, Bil98, BP01, BBD^{+}09, BPT06, BB10, BL03, BC05, Boe01, BGPW05, BG10b, BBD97, BGP00, BCP94, Boo91, Bor93, BGR03, BC08, BAL88, Bos00b]. Italy [BEGP04, BEG04, BSB^{+}99, BDV95, BZ09, BGR08, BW09, BSS00, BPSW06a, BPSW06b, COBV11, CFI06, CD10, CLM03, CV04a, CC05, CT03, CNP^{+}08, CGS01, CVLL12, CW94a, CFGN11, CFH^{+}12, CR01a, aCF06, CMMR12, CC91, CM03b, Chi03, CNY95, CH96, CW06, CGP03, CB06, CS12a, Coe11, CGK^{+}12, CCD^{+}12a, CCD^{+}12b, CF09b, CGS02, CSAH09, CDSS07, CR73a, CR73b, CGGO03, CLS07b, CLP05, CAdB01, CM09c, CR04b, CF99, Cor02, CCSW12, CP05, CDFP11, Cou93b, Cra12, CPP07, CDM04, DGT10b, Dal92, Dam99b, DLV04, DSLT10, DFUR11, DR05, DDFM05, DF09, DD95, DFM04, DPZ12, DY06a, De 95, DGMS97, DG09a, DEG11, Del97a, Del97b, DP11, Dem07, DFG^{+}12, DGLT12, Di 97, DSV01, DCC^{+}11, DBD^{+}11, DPP09, DKD05, DPT12, DPP95, DO09, DO11, DFF06]. Italy [DFT12, DLO03, DPNM12, Dro09, DRS08, EL07, EEK^{+}87, EKLM87, Ehr04, ÉEMS04, Esp01, ERMS06, EFZKM07, EHMM09, EEM^{+}11, Eta05, FSA^{+}07, FBKS03, FPP11, FM97, FL08, FMZ05, FFL^{+}12, Fes10, FdCLN^{+}11, Fic03, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, FBV11, FB09, Fle02, FSV09, FP08, FJ10, FYM12, FS09b, FC93, FL11, FdS12, FT94, Fuc10, FKL08, FMC12a, FMC12b, FMC12c, GE08, Gan99, GD10, GT03, GML^{+}08b, GML^{+}08c, GMP06, GG02a, GS09a, Gia04, GBC^{+}08, GM11a, Gia89, GG98, Gir99, GOW03, GT04, GS05, GLN01, GS95, GW06a, GS10, GL08, GCC^{+}02b, GCC^{+}02a, GSS11, GVT^{+}11, GI03, GG02b, GYD^{+}06, HP99a, Ham08, HKI08, HS95b, HL10, HSS96, HM01, HH02, HMPZ09, Hus01, JLV02, JS08, Jon01]. Italy [KM08a, KL92, KR00b, KGG^{+}11, KZ01, Knu99a, KV08, KKL12, KMVY06, Küh07, KLT09, LGK^{+}04c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, Lag05, LM93, LM00, LS10a, LPOA^{+}06, Lee11, Len97, Leo04, Leo09, Lev98, LMP04, Lio04, LM03, LPR08, LMT10, MP98, MF11a, MF11b, MST12, MBW08, MS03a, MSS11, MM08, MY01, MD04, MT02a, MSY08, MB01, Mar03, MP10b, MWZ10, MMP07, MTV09, MPT10, MPPJ^{+}05, ML11, MSW05, MP03b, MTS^{+}03, MT^{+}03, MRSV07, MH11, MM82, MN12a, MN12b, MSH08, MMMO96, Mil01, M^{+}06, Mio90, MR97b, MT99b, MR02, MS96c, MSB07, Mor89a, Mor06, MK03b, MO05b, MBC^{+}97, NDFMS97, NR99, Nes95, NCBT04, NWHR06, NR06, NBC^{+}06, NVSB^{+}07, NSL09, NSS03, OVG^{+}08, OFR09, OWPY03, ODV08, PSP^{+}02, PAV^{+}10, PGM98, Pal01, Par98, Pat02, PdRM^{+}12]. Italy [PFD^{+}10, Paz97, PZ96, Pel11, PH97, PH11b, PT98, Per11c, PT97, PPS03, PB09, PH12, PS11a, PRG11, PZ07b, Pri03a, Pri03b, PQ05, Pri06, PW08, Pue07, RKS89, RTA95, RRR01, RtTMP10, RF99a, RL11a, RK02, RKW04, RV05, Rol97, RF12, RT09, SBV^{+}11, ŞSW07, SS11b, SÇS12, SYMA09, San92, San01, SKR11, STT11, Sch03a, SCM88, Seb03, SVS08, SC09a, SG09, SMP07, SBD03, SFN^{+}11, SPO11, SEN08, SC09b, Squ09, SL03, SWMM06, SGD09, Stü09, SA04, Tah01, TBE08, TC95, TFE^{+}09, TBC07, TBG05, TN09, TP94, Tor93, TEB09, TV87, Tuc01, TAS05, VC03, Ver08, VD12, VSTN06, Wil01a, WSF08, WKB04, Yah11, YRMD05, Zil87, dSG05, dlBP08, vHC08, vF91, vLS91, vdHCD07, vdL04, vzGIK08]. Itatiaia [V^{+}06]. ITBAM [BKLP11, BKLR12, KLP10]. ITC20 [MDY07]. ITEMS [YYH^{+}12]. Iterated [Geu98]. iterative [Fre87, Lin01]. ITG [MCER10, Sch12, BB95]. Ithaca [AH94, Ant97, CR05a, FW07, Pin95, Van02, Zip92, WF77]. ITIB [PK12]. ITM [NMP^{+}04]. ITMAS [DHtM^{+}12]. ITP [BF12, KP10a, vEGSW11]. ITRUST [HIS05, JPD04, SWMM06, NT03]. ITS’98 [Goe98]. Ittingen [KWT03]. ITWP [MA05c]. IUKM [THL11]. IV [AG07, BE06, CR05a, DL04b, EL07, FKKB^{+}05, GCL^{+}06, GTM09a, HKW^{+}11b, HMBD09, KHSM^{+}08, PSP^{+}02, PVP05, SYY^{+}08, AvASD06c, Ant97, Bos00a, BvASD04d, BD03, Bur90, CGJ^{+}11d, DMP10d, EV96, F^{+}12b, FTZ08d, GGK^{+}06d, GGK^{+}05d, GMO03, HSNJ02c, HZLL12d, Jac07a, Jac09b, Jac11d, JDK10, KKS11f, KK00, KDH^{+}11d, LGK^{+}04d, LBP06d, MGI^{+}11d, MGM^{+}12d, Ngu11b, PM04d, PCM^{+}10, PS05b, PCE06, RA07b, SJHJ10d, SvADS07d, Shi09, SRW98, SAB^{+}03d, Spa03, Ste11b, Ste11c, SBK01, TGM^{+}10d, dLGR07]. IVA [ABB^{+}10a, GYA^{+}06, NNPW12, PGA^{+}05, PMA^{+}07, PLI08, RKNV09, VKMT11, DAB01, RABR03]. IVIC [ZRP^{+}09, ZRP^{+}11a, ZRP^{+}11b]. IVR [PPCS12]. IVR-XXIV [CPSA10]. IW [vdL01]. IW-SAPF-3 [vdL01]. IWAAL [BHV11]. IWACA [Ste94]. IWAN [HDLM09, MCSY07, MNW01, STTP02, WSS04, Yas00]. IWAN’99 [Cov99]. IWANN [BHV11, CPS05, CSPC09, CRJ11a, CRJ11b, HC11, MMMD97, ORB^{+}09, SPCG07, MCP93, MP01a, MP01b, MA03b, MA03c, Pri91]. IWBRS [LST^{+}05]. IWCC [CLL^{+}09, CGL^{+}11, GNTF02]. IWCF [GFV09, SFS11, SF08c]. IWCIA [ABB^{+}11, BBA12, BBH08, REF^{+}06, WB09, KŽ04]. IWCM [JMBS07]. IWCMQ [JP03, OYY03]. IWDC [DSDS04, PKKG06, Pal01, DB02, DD03]. IWDE [OD12b]. IWDM [ABRZ06, Kru08, MBG12, MOFM10, BF89]. IWDW [BCKK05, CKL05, HSKB09, KKH09, KSB11, SJ06, SKK08, SKPG12, KCR04, PK03]. IWGD [SPÖ^{+}10]. IWHDM [IHX^{+}08]. IWIC [IFV07]. IWICPAS [ZJL06]. IWINAC [FSdlPT11a, FSdlPT11b, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b]. IWKDEWL [LCY^{+}11]. IWLCS [BBMB^{+}08, BBD^{+}10, KLT^{+}07, LSW01, LSW02, LSW03]. IWMM [LOS05, Bak95, BC92]. IWOCA [FKM09, IS11a, IS11b]. IWOMP [CZG^{+}08, CGKM11, CMMR12, EdS08, MCdS^{+}08, MdSC09, SHM^{+}10]. IWPEC [BL06b, CF09a, GN08, DFD04]. IWQoS [dMB05, JSW03, WHS01]. IWSAS [LRS03, RSL01]. IWSDS [LMMN10]. IWSEC [EKS10, HY12, IN11, MF08, MKR07, TM09, YSR^{+}06]. IWSM [ABD^{+}09, CGBDA08, DBB^{+}08c, DA01]. IWSM-Mensura [CGBDA08]. IWSN [BGG^{+}12, SLL^{+}06]. IWSOS [BG11b, HS08c, HK07, KH12, SH09, dMS06]. IWSSA [MTH08, MHD09]. IWVF4 [CAdB01]. IWWERT [AP93, Sch92]. IX [APV09, CAB10, DS09a, FST09, GTA10, HMT10b, LGS09, RBB^{+}06, SSVW12, BJNT06, LCH^{+}12, PSR08, Pri08, SAF^{+}07]. Izmir [HWW^{+}10, Sav06, Yak00, Yak02, Yak04, YN06]. IZS [Pla96].

J [ACÇ^{+}12, HNR05]. Jack [JRR03]. Jacques [Nac12]. Jadwisin [Bli75]. Jahrestagung [Bra73, Müh75, Sie75]. Jaipur [CD00b]. Jalisco [BCE09]. Jamaica [BW05b, Bru06b, SV05]. Ján [VBCBS05, MvOvRdV05]. Janeiro [BPAB07a, BPAB07b, BC09b, DPC^{+}07, ERK07, ES03, FMS10, LdQ07, LSHB07, SMK01, ZZTG07]. January [AK97, AYV^{+}11, AFKL04, AN09, BS11c, BH10, BCL03, BB99, BC10a, Ber02b, Ber03, BB01b, BFG^{+}12, Bih97, Bis88, Bjo80, BB10, BTZ^{+}10, BDDM12, BKM06, BS01c, CLM03, CP10b, CA08, ČGH^{+}11, CCD^{+}06a, CCD^{+}06b, Che03b, CP98b, CP98c, Cho02a, Cho02b, CK06c, CH99, Coe11, CP07, Cor02, Cou05, CK82, CKLM09, DW03a, DKS^{+}07, Des02, DH94, Dik04, DL04b, DSS^{+}07, EL04, EN06, EHSH92, FL00, Gae96, GWK09, GKB^{+}08, GS09b, GM99, GV95, Gup99, Hae99, HS12b, Han07, Hen06, HC05, HHLR08, HKV10, HNPP07, HW08b, HSMPA09, IMW04, IMK06, IZ00, IFV07, JS11a, Jon80, JMO09, JMP12, JK06, KS09a, KPSW10, KWW04, KM08b, Kim05a, KN83, KR02c, KH09, KZ10a, KR12, KMMP12, LV07a, LTL^{+}11a, LTL^{+}11b, LLM^{+}08, LN03a, LPZ08, Mar11c, MB01, Meh85, MVN86, MS02c]. January [NNS06a, NNS06b, NKM^{+}09, PR07, PS98a, PFD^{+}10, PC00, RS09a, RUL04, RRU05, RCJ^{+}08, RL11b, RZ12, SNE08, SE94, SMH^{+}12, SS06a, SP90, SEO^{+}09, SLL^{+}06, Sio10, SCD^{+}10, SL99a, SLL01, SL03, SDK^{+}08, STWS07, Stü09, TLF01, THR03, Tsu08, VPBŠ03, VP96, Vau05, VFC08, Ver08, VBCBS05, Vyt91, WHL09, Wat95, WTP^{+}06, Wri03, WVDS05, ZJ06, Zha07, ZLS^{+}06, Zuc03, vLIvdH^{+}07, vLMP^{+}10, vdBV01]. Japan [AI97, ANS04a, ANS04b, ANS04c, AKU01, AK03, ABK05, AKT06, ACK^{+}07, AB04, AT07, AT12, AT02b, AM98, AF99b, AM00, AHMT08, Asa96, Asa99, AiNOW11, AH07, BE06, Bar03, BIRT05, Bha00, Bha02, Bha05, Bha07, BB03, BR96a, BKV10, BI04, BDDS07, BKNS06, BGY96, BHJY07, BFK^{+}07, CDP02, CHM^{+}10, CWiM^{+}09, CK06c, CQSS09, CKOB11, CTS07, DMO98, DvEF06, DK02a, DK02b, DSL08, Dos93, EKS10, EYX^{+}05, FS07a, FvdHJT10, Fur95, FU96, FTF88, FTF89, FTF91, FM96, FMY04, GMO09, GH08, GJT03, Gel11b, Gel11a, GMV09a, Gie05, GHI^{+}12, Gom01, GAY86, GM06b, GHH^{+}06a, GTY03, GY03, HM94, HO03, HW06a, HI93, HY12, HKI^{+}09, HM12b, HML97, HRA02, HST07, Iba92, IKO03, ICSY07, IMK06, IRM93, IZ98, IZ99, IST07, IK12, IGN05, IFV07, IDMY08a]. Japan [IDMY08b, IS02, IH90, IM91, IY95, IT03, IKKU08, IT08, IN11, JSS08, Jan93, JSZ95, JDMB08, JMO10, KS04a, KKKS11, KUA00, KM08b, KSB10, KMSB11, hKA10, KKKN05, KILW10a, KILW10b, KP01a, KHSM^{+}08, KS05a, KU01, KMMP12, KJS01, Kur10, KL02, LJS08, wLN99, LSL^{+}09, LMMN10, LO98, LM97, Liu01, LMA08, Luo12, MB82, MF01, MMT97, MKT98, Mat02, Mat09, MF08, MPR07, MPvH04, MS99b, MKR07, MNTY05, NMM10, Nak92, NcC99, NYHO07, NMPM08, Nip98, NY93, NK99, NS81, NABK12, NVSB^{+}07, ODP^{+}07, OT11, Oka03, OS07, OYM07, Oka00, OBS12, OA07, OS03a, OBM11, OKdQ09, PZ07a, Pol97, Pol99b, PLI08, PT11, RS99a, Sad13, Sah13, SKS^{+}06, Sak91, SFS11, SHN07, SHM^{+}10, SNE08, SINK08, ST12a, SKSA12, SILS05, SHRM06, SHYEF08, SHUH08, TOOO07, TM09, TPB79]. Japan [TvdBI02, TFK11, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, TLC00, TBF^{+}09, TO08, Tok07, TNM05, TNY07, TNI09, TYNM05, UWB^{+}06, UMN^{+}08, Urz05, VJKT00, VJAA03, Wad86, Wad87, WHL09, WSS04, WSTI07, WSTI08, WY99, WZ09, Wei97, WPM07, WVDS05, XSA^{+}12a, XSA^{+}12b, YKKZ07a, YKKZ07b, Yam08, YGGJ04, YAL^{+}05, YYI^{+}09, Yas00, Yi05, YI91, YM01, YMY^{+}10, YSR^{+}06, ZSO99, Zho01, ZRTS03, ZJS^{+}02, vLWH^{+}00, vdHI99, vdHIP11, BGY96]. Japanese [AKU01, AK03, AKT06, CKOS09, DEM96, KUA00, AS02b]. JapTAL [IK12]. Jasná [Rov98]. Java [GAR03, GAR04, GRR05, AF99a, AJ01a, DHM99]. JavaCard [AJ01a]. JBI [FN12a]. JCDCG [AKT06, AKU01, AK03]. JCDCG’98 [KUA00]. Jean [CLKK07, Mah04, Nac12]. Jean-Jacques [Nac12]. Jean-Louis [Mah04]. Jean-Pierre [CLKK07]. Jefferson [DPP98, Fre98]. Jeju [ACP05, Bai05, BL10b, BBK^{+}11, CY03, CCP10a, CCP10b, CSY09, CY11b, CM04a, DWSK05, HK05c, HTMC09, JY12, Kim05b, Kim05a, KYL07, hKhLKŚ10, hKAS^{+}11, Lee04, LLWL04, LYMM05, LYY07, hLhKcFŚ09, LY12, LY05, SJ06, SKY06, SKCK07, SHŚ^{+}07]. JELIA [APO96, FvdHKL06, HLW08, JN10, OAdGBP00, PW92a, dCHM12, vE91, AL04, Fle02, MPP94]. JELIA’98 [DdCF98]. Jena [DNS09, GHP^{+}11, KK09a]. Jersey [FC98, TT95]. Jerusalem [AS94, AL01, BB99, BD97a, DSM94, DL95a, GSLM12]. Jhongli [IAC06]. JiangSu [CL04]. Jinhua [ALZ11, LWM^{+}08, LZS^{+}08]. JIST [PCK^{+}12]. Jiuzhaigou [CTYG10, SPÖ^{+}10]. JML [Mos97]. JMLC [LS06a, BS03b, GW00]. Job [FR95a, FR96, FR97, FR98, FR00, FR01a, FRS02, FRS03, FRS05, FFRS05, FS07d, FS08, FS09b, FS10, RF99b]. JOBIM [GS01]. Joensuu [KPK05, RRM06]. Joerg [HS05b]. Johan [OKL04]. Johannesburg [CP11, LW12]. John [BCGY09, BL11, NH03, YDW08]. Johnstown [BCR03]. Joint [AL11a, AM91b, AM91a, AI06, ABG^{+}06, Ama08a, AS98, AFR^{+}04, ABD08, AL11b, ALZ11, ALL^{+}10, BBE01, BH06b, BCS09, BB90, Bar11, BR04, BPR05, BD97b, Bir12, BMO02, Bod05, BS03b, BCRT92, BKL^{+}12c, BD11, BV94, Bur90, CAD^{+}02, CBC^{+}02, Cas09a, CR01a, Cer05, CW09, Cle99, De 07b, Deg03, DV09, DG12, DO89a, DO89b, DLW^{+}07, Dro08, Due04, DL07, Ehr85a, Ehr85b, EEK^{+}87, EKLM87, EIRW06, EM10, FPFR05, FI08, FOGE^{+}09, Fin99, FK12a, Foo99, FCD^{+}04, FS06, Gab97, GH03, GJ93, GO11, GR12, GH10, GHI^{+}12, Gor10, GLN01, GS00b, GMS12, GH07b, GHP^{+}08, Gup10, GW00, HZ05, HWW^{+}10, Han98, HZ10, Hed03, Hen08, HS02, HP06, HSP^{+}12, HCF06, Hof11, HM01, HSL^{+}99, Hor02, Hus01, IB07, IZP09, Jäh99, JP04a, KS02c, KMS^{+}11, hKA10]. Joint [Kno11, KLLP12, Kos98, KP09, KO07b, Le 02, LLPH09, LFP^{+}09, Li11a, LS06a, dVLKR^{+}08, LSU07, Mai00, MY01, MS02a, MP12, MS00c, Mos97, MNS95, MSY00, MZ06b, NL09, NE02, O’B12, Ong10, OT07, PCK^{+}12, PdRM^{+}12, Pez03, PPBL12, RR08, RT10, Sag05, San01, Sas05, Sat97, ShK06, Sch04b, Sei07, Sei12, Ses06, SCR06, Smo00, SLSW12, SS08b, SCB01, Swi99, TC95, TNN^{+}01, Tiu00, TGM97, Wal04b, WR11, WSN^{+}06, Wat00, WY04, WK02, WMS04, Wil01a, WZ08, YKF^{+}06, ZJ05, dA09, dLZ12, dMS09, dRWK^{+}10, ABK05, Ami98, Apt00, BC93, DWSK05, DLT05, DG01, DN93, EO94, Est98, FIAP01, Gor03, HHM97, HOD96, Kay03, LY04, MO92a, MZF94, Niv98, O’S03b, RTA95, SSW92, Ste98, STLK05, Tho99]. jointly [HAB98, JS97, LdRV94, NL99, PGM98, ŠPTM00, SBH^{+}04]. Jolla [CF00a, Ran93, VS05]. Jones [MSS02]. Josas [FPMW06]. Jose [BFMS08, LG94, Poi06, Pre02, Seh97, ZSM^{+}09, CGS07, KPR07]. Joseph [FJM06]. Jouannaud [CLKK07]. Journal [SMA03, Spa03, SBJ^{+}05, SAC^{+}06, SACM06, Spa06, SAF^{+}07, Spa08, SPT^{+}08, Spa09, SZS09, SD09, Spa11, ZS05a]. Jouy [FPMW06]. Jouy-en-Josas [FPMW06]. JSAI [HKI^{+}09, NMM10, OBS12, OBM11, SHN07, SINK08, WSN^{+}06, WSTI07, TNN^{+}01]. JSAI-isAI [NMM10, OBS12, OBM11]. JSME [SLF91]. JSPS [FMY04, Oka03]. JSSPP [FRS05, FFRS05, FS07d, FS08, FS09b, FS10, FR00, FR01a, FRS02, FRS03]. JSSPP’99 [RF99b]. JSSST [FM96, NY93]. JTRES [MT^{+}03]. Juan [BP04a, BTBP^{+}08, CMM93, Rol99, RF99b, AF02, DC00]. Juan-les-Pins [BP04a, BTBP^{+}08, DC00]. Juli [BI73]. Julians [BFF^{+}10]. July [ÅBE12, AJS05, AS94, AGK^{+}10, ADG^{+}08a, ADG^{+}08b, AHS11a, AHS11b, ABH12, AN97, AW01, Akm01, AN95, AG98, AMSM^{+}09a, AMSM^{+}09b, AV08, Alp11, ABB99, AH96, AP04b, Ama08b, ASLM09, Ame91a, AL01, AHK07, APHA11, AS98, AJ11, AH97a, AT02b, AFR^{+}04, ABM05, AK12, AF06, ACJT07, AL12, AI98, AHMT08, Asa99, AH10, Atl08, APM08, AB78, ADCR89, ACR^{+}08, ACD^{+}10, Ayk07a, Ayk07b, Ayk09, Baa03, BBMB^{+}08, BBD^{+}10, Bac00, BO00a, BLPW03, BHL^{+}03, BEOT06, BFH07, BA07, BR04, BGP11, BPR05, BS02a, BSN06, BW10a, BS76, BGV01, BBP^{+}12a, BBP^{+}12b, BD12, BE87, BEM91, BBLT06, BH06c, BAHH07, BE08, BEP04, BCKR10, BR06, BDW08, BF09, BDSW09, BFS92, BK80, BM10, BVAT12, BG03b, BdBK03, BMW93, BBKM11, BBP93, BBZ00, BBZ01, BSS11, Bla05, BGU95, BPD06]. July [BM02a, BJ01a, BLvdH10, BDDS05, BDDS06, BC08, BG08, BK04a, Bos00a, BE07, BM09a, BM94, BMYZ91, BMYZ95, BMCCM11, BDN08, Bow94, BCD00, BG05, BN09b, BKLS11, Bra09a, BTT12, Bra85, BL06c, BGHS05, BHK01, BL02b, BEFP08, BRW85, BZ03, BSH06, BM11b, Buc90, BPSW06a, BPSW06b, Bun94, BL06d, BBC04, BKO03, BFKO06, CIM^{+}05, Cal86, CBC^{+}02, CDV03, CCJS07, CRL01, CPFD^{+}03a, CPFD^{+}03b, Car03, CS05a, CDCW09, CW94a, CL01, Ceb08, aCF06, CZ05, CM03a, CWZ^{+}11a, Che10, CTYG10, CW00, CS05b, CNY95, CVV12, Coh92, Coh94, CGM95, Coi96, Coi99, CH91, CSAH09, CBF01, CMP07, CK07, CFP^{+}09, Cou93a, Cou01, Cov99, CPR12, CRYS09, CFL10, CRW^{+}12, Cue03, CBCGA12, Cur09, CFB11, CF00b, CT99, CMPW12a, CMPW12b, DFL^{+}10, Dai07, DL06a]. July [DH07a, Dam99b, DF10, DDE04, DFUR11, DM01b, DLT05, DG07a, DCFR12, DdQ10, DG96, DB98a, DR05, De 10a, DW12, DGMS97, DSS03, DFP^{+}12, DS01a, DB98b, DL95a, DJS06, DJNR06, DKLT05, DEM96, Dia83, DKLS04, DF96, DVM^{+}11, DN09, Dig04, DvEF06, DMRS06, DFN97, DB07, Dol09, DOSY05, DP12, DTF98, Dro09, Du00, Dub09, Duf07, Duf09, Duf11, Dyc00, ENP02, EJM10, EF02, EV01, EM05, EH08, ELNM08, ELN^{+}10, Emb98, ES00a, Eng81, EKO^{+}04, EL02a, Erl06, Ern07, EF03, ER05, EKM06, Eug09, EK81b, EH12, FBKS03, FC98, FPRAW04, FRA05a, Fed01, FRS02, Fek08, FGB11, FLVS11, FLMG10, FKM09, FK02, FHH11, FHW12, FDL10, Fit84, FHT05, FB09, FG06, FK12b, FG10, FLL03, FPRS06, FLR11, FK95b, Fuc98, FG95, FTF91, GO94]. July [GU95, GHM05, Gan96, Gan99, GB99, GL10b, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, Gia89, Gib89, GW09, GH10, GK90, GH06, GMO03, GHSW04, GLM^{+}08, GOW03, GK95, GM06a, God05b, GW11a, GQ11, GHC^{+}06, GS88, GS10, GP08b, GL92, GJS08, GW06b, GM11b, GW93, GV09, Gud08, GI03, GNR94, GPP09, GY03, GM08b, GPL99, Had03, HK04, Hal94, HP99a, HENR03, Hal00, HPTV05, HS07, HP99b, HV03, HMT10b, Han09, HS06, HKL07, Har07, Har09, Har11a, HMTH10, HMT99, HS08b, HMW^{+}08, HW01, HBN06, HS02, HPP06, Heu98, Heu00, HW09, HKLP12, HdQ08, HŽ07, HB11, HKP11, HS98b, HPRS06, HRRY08, HNP12, Hor06, HV98, HJBF12, HMJG12, HMBD09, HKH10, HS03, HP99c, ID03, IR08, IPSK04, IS11a, IS11b, IMY^{+}07, Jac07c, Jac07d, Jac07b, Jac07a]. July [Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, JNS05a, JLY03, JR98, JL00, Jen01, JCC^{+}12, JV06, Jul98, JK05a, Kah79, KC09, KJR05, KA12, KJ88, KvETvdH10b, KvETvdH10a, KS12a, KL88, KL96, Kar09, KN91, KL07b, KP10a, Ker01, Kim01b, KS06, KK98, KR01, KS02d, KW98, KSW99, KK00, KFW04, Koh06a, KBHSW06, KCMO11, Kot86, Koz04, KJ10, KP15, Kru08, KSTP99, Kuc11, Kui92, KRS02, Kul09, Kur09, Kur11, KMR12, KCK08, KMKP11, LSJ06, LSV06, LVS08, LSW00, LSW02, LS92, LSW98, Lau01, LLCL05, LP97, LCHS08, LOST04, LOTY05, LTÅ^{+}11, LdQ07, LLC^{+}05, LS88, LNE06, LM08a, LSE02, LV06, LWF04, LWD05, LVW05, Li11b, LPR12, Lin07, LMP04, LKC93, LJ00, LNS11, LLL06, Liu11, LK04, Llo00, Loe74, Lor01a, Lor01b, LG05, LNSS11, LMŠT01, LSW08, LR05b]. July [LM08b, MZ07a, Mac12, MS12a, Mad92, Mag98, MBG12, Mak86, ML04, MSP^{+}09, Man09, MMBND07, Mar11a, Mar11b, MSH85, MST92, MVMP^{+}04, MTCOBH11, MJR04, MMDN12, MMP07, MS03b, Mau79, ML11, MW97, McC97, MW09, MS96b, MMO^{+}11a, MMO^{+}11b, MR08, MT89, MM96, Mez11, MKZ02, MKZB04, MKZK06, MKZK08, MKZK10a, MKPZ12a, MKPZ12b, MR07b, MHK05, MV12b, MS08a, MOT12, MKB^{+}09, Mol95, MTW05, MAA91, Mor89a, MR12c, MK03b, MSS05a, MBA05, MK96, MK12, MSY00, MSS08, MTNL12, MSR99, MZ06a, MA07c, MYS10, NR99, NKPH08, NT92, NM94, Nes99, Ngo09, NS82, NZ03, Nie93, NRWI09, Nie05, NBK04, NR04, NR06, NP11b, NG10, Nos88, NvRSS11, NU92, ONC^{+}11, OO12, OGM05, OLL10, Oli07, OPT01, OSvL01, Ott87, OSRS07, ÖRH12]. July [OZ09, OZ11, PSP^{+}02, PAV^{+}10, Pag98, Pal00b, PZK07, PH11a, Par84, PMMR06, Pat90, PC99, PMH09, Paz97, PZ96, PW09, PLC11, PS02b, PF06, PF08, PF10, PFM12, Per01, PR03b, Per05, PI05, Per06, Per07a, PS07a, Per07b, Per08, PS08a, Per09a, Per09b, Per10a, Per12a, Per12b, PN93, Pfe94, Pfe07, PGD07, PV04, PT99, PLG06, PP11, Pol01, PvS07, PISW05, PLMV12, PMS08, PPM09, PCA02, PPH07, Pue07, PT06b, QBA01, RCC12, RMR05, RFZ03, RMS04, Rau11, Rea01, RR03, Rei88, RG03, RW06, Ria08, Ria09, RtTMP10, RtTM12, RB91, Rob11, Rol97, RBSL06, RDK09, SP09, SNS03, ŞS05a, SMD04, SL12, SS77, SS09a, SS11c, SMB03, SHØ06, Sch10a, ST11, SS01, SS94, SCHS07, SR07, SEG09, SF11, SV97, SNRG11, Sch07, SSN08]. July [SLZH07, SVA05, Sex09, STS04, SPÖ^{+}10, SXC04, SILS05, Shu07, Shu09, Shu11a, Shu11b, SCG98, SBD03, SA06, Sie86, SRW98, SA12, Sma97, SS07c, SS07b, SS09c, SS11d, SPO11, SC09b, SMCM07, SH10, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, Sti90, SS10, Stu02, SDM05, SC07, SNSR06, SMS12, SH11, Tah09, TOOO07, TN03, TS10, Tho06, TGH06, Tis02, TWN92, TP94, TNM05, TNYL11, TCJ10, TG10, Tre09, TMN09, TOL08, TtHVS06, Uus06, VGP08, VTK^{+}12, VY93, VPM97, VM01, VWH06, VBH07, Ver00a, Ver00b, Vid12, VT97, Vit08, Vit10, VZ99, Vor92a, Vor93, Vor02, Vor08, Wad86, Wal84, WYL01, WPV04, WPSY06, WYP12a, WYP12b, WZ03, WW03, WMR00, WLP^{+}09, WPM05, WPM06, WTM^{+}02, WP10]. July [WM04, WN96, WPD04, Wol95, XYM^{+}07, YBM87, YJKSW10, YP01, YY01, ZSSM03, Zam08, ZL98, ZJ04a, ZPTI09, ZHLW12, ZS05b, dGEK11, dMLG03, ter97, vEHD05, vWN99, vLdB80, vL88, vdMvdT08, vdMS11, vP93, von96, vzGIK08]. July14 [MKZK10b]. July14-16 [MKZK10b]. June [AK10, AG01, AvLR91a, AvLR91b, APFV09, AK07, AGB^{+}11, AM10, AJNJ12, AMS06, ASS10, AAA^{+}11, AS99, AB12, ABF^{+}07, AVP07, AM97, ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12, AM91c, ABDW10, AKM10a, AKM10b, ABMS06, AE05, AD06, Amb01, AP10, And96, ABJ02, ADMH04, ADM^{+}06, ADM^{+}08, AGS^{+}11b, AGS^{+}11a, AJ11, AMT03, AMNT06, Apo93, AH97a, ACP05, AFR^{+}04, AMPT09, AZ11, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, AN07, AB03, APP12, Asp98, ASAM07, AAR05, ABRZ06, AL06, AHSC11, AM05, ADP11, ADS09, AB97, ABFR07, Baa07, BVO10, BW05a, BLPW03, BYCC03, BCBF08, BL92, BB07, BR12a, BW11, BSZ12, BSW95a, BDGM05, BPCS06, BHJKPG06, BG93a, BM05b, BdB08, BRS^{+}07, BMH05, BG91b, BHHK08, BS06a, BM11a, BBLT06, BDL08, BL08, BR92, BGKY08, BKR09, BB05a, BP09a]. June [BC96a, BS92, BMP90, BCG^{+}11, BH07, BDZ08, BPZ09, BI11, BCP12, BBKW03, Ber00, BHCL87, BM99b, BN04, BVAT12, BS01a, BG03b, Bil97, BdBK03, BR96a, BMW93, Bis02, BV11, BBKM11, BBRM98, BBP93, BBP96, Bla03a, BvDMST11, BS10, Bla80, BS02b, Bli75, BRST08, BRW93, BR12b, Bod03, BRU06a, BU01, BHU03, BFSN10, BCC^{+}12, BG10b, BDK10b, BO00b, BKS05, BJ12a, BJ07, BF89, BG07a, BM98, BM09a, BIMW12, BZ08a, BKL^{+}12c, BKM01, BGKS04, BW00, BL01, BKM07, BAH09, BHV11, BGHS05, BGY96, BEFP08, BP12b, BJ12b, BtHRBB12, BD11, BS12, BCR03, BG07b, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BvADS08c, BvADS08a, BvADS08b, BBL87, Bue04, Buh98, BBB06, Bun94, BN98a, BN98b, BDLMM12, BFOY04, BS11f, BFK^{+}07, CPS05, CSPC09, CRJ11a, CRJ11b, CV11]. June [yCW96, COBV11, CHM^{+}10, CKPR11, CK12a, CK08, CK10a, CK10b, CK12b, CK12c, CP06a, CS07, COMTLB12, Cas01, CGP02, CCPP12, CZ05, CT07, CWC^{+}07, CZG^{+}08, CGKM11, CMMR12, CLL^{+}09, CGL^{+}11, Che03a, CR04a, CMZL04, CChL^{+}10, CP06b, CHCA09, CKH10, CCBW11, CHAL04, Chr00b, CHA03b, Chu03, CW01, CW06, CD05, CS12a, CA10, CVV12, CBT07, CK01, Com97, CDSS07, CR01b, CDW86, CLT05, CLS07b, CDL12, CDNS00, CWO^{+}09, CRS10, CBW99, CM09c, CR04b, CDFP11, CM04b, Cou93a, Cou03, Cov99, Cow77, CPP07, CR05b, CG94, CMQ96, D’H10b, DFL^{+}10, Dam97, DM06, DG06a, Dav89, DM01a, DDHP05, DPC^{+}07, DKC10, DSB10, DMR01, DR11, DD95, DS09b, DPB^{+}04a, DPB^{+}04b, DFP^{+}12, Dem07, DT89, DHX09, DV07a, DFG^{+}12, DGLT12, Des86, DS98, DPW04, DLN10]. June [DPP09, DN09, DMWW88, DJJ12, Dil94, DGN01, Dit01, DPP95, Dji89, DFT04, DT06a, DLW^{+}07, DW94, DMW06, DPNM12, DHP09, DP06, DG97, DNS95, DF99, EB04, EM03, EKO08, ES10, EKP^{+}07, EM06, EK10, ELN^{+}10, EML^{+}12, ES03, EW91, EP07, Esc10, EV07, EL02b, ES92, EH12, FHJ^{+}02, FRS08, FPFR05, FLS06, FK10, FR01a, FRS03, FRS05, FFRS05, FR11, FKK09, FL08, FSdlPT11a, FSdlPT11b, Fer08, FLMG10, FMZ05, FS07c, FKM09, FV09, FW07, Fis08, FX08, Fle99, FHM82, Fom06, FLUV02, FJ10, FS07d, FS08, Fra00, FKBP11, FW09, FKM07, FRSH05, FL11, FGM92, FGL96, FKM08, FKS07, FT94, Fro02, Fuc10, FCDB07, FTF88, GO96, Gab97, Gab13, GGD09, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GVCA07, GB99, GY08, Gau93]. June [GFV00, GMV03, GB97, GKS^{+}07, GL10b, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, GS00a, GM11a, GAA06, GN12c, Gie10, GR12, Gil10, GLM^{+}08, GH09, GSH^{+}04, GSS09, GW11a, GZ09, GGHS07, GSLM12, GPE05, GdlP99, GLN01, GLS05, GZSC07, GW06a, GHC^{+}06, GH12b, GS88, GR06b, GMS12, GL03, GHH06b, GF12b, Gru97, GRP10, GDGB12, Gue99, GS94, GW11b, HNH^{+}09, HRZK02, HP12, HO03, HG02, HRdV11, HV03, Han81, HB97, HK12, HH10, HMT99, HZ10, HKI^{+}09, HKP10, HMW^{+}08, HJ12, HA02, HM02, HC11, HHW01, HRSS08, HM96, HCF06, Hof03b, HCR06, HDGK11a, HDGK11b, HRM^{+}10, HMMW10, HSL^{+}99, HH02, HT06b, HPSZ08, HMPZ09, HPP88, HS98c, HNW04, HV98, HW08a, HC09, HH03, HP89, cHPkC^{+}07, HS99, HKH10, ID06, ILR95, IAC06, Ima99, IR07, IL01]. June [IKY05, IS12a, IS12b, IH90, IKKU08, JYZ04, JO99, Jay04, JS08, JNHJ10a, JNHJ10b, JSW03, Jen92, Jeu98, JKN^{+}12, JDAW12, JRPP07, Joh01, JP11a, Jón12a, Jón12b, JS02, JJF03, JP12b, JK05b, Kåg98, KEDW07, KMP84, KPK05, KKGK11, KC11a, KC11b, KL07a, KSS08, KO91, Kap10, Kap92, KGL04, Kat92, KY07a, KS09b, KWC^{+}12, KKMW07, KLT^{+}12, Kay03, KLM^{+}07, KMF06, KD05, hKA10, Kir93, KR00b, KN03, Kla12, KD99, KY07b, KS93, Knu01, KvG98, KK11d, KHSM^{+}08, KZ97, KS05a, KC02, KFMM06, KV08, KK09b, KY11, KS05b, KHI06, KMMT12, KGS08, KC84, KS04c, KKL12, KLFK10, Kra05, KSHM09, KP11, KOS07, Kro00, KF74, KPRS07, KRSR11, KSTP99, Kuc̆02, KU09, KSGT10, KR04b, KV11, Kul09, KA05, KMKP11, KŚHM11, LH11, LM98a]. June [LM06a, LO11, LPOA^{+}06, LZ08, LSgL^{+}06, LLPH09, LBM^{+}10, LPH09, LVW05, LP10a, LMWY04, LMW06, LMW08, LMW10, LMW12, LS89, LR94, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, Liu11, LABM^{+}12, LS04b, LMT10, LPA00, LSF07, LT11b, LP00, Lov82, LNSS11, LZ00, LS06c, LHH11, Mac12, Mad92, Mag03, Mag02, MMM01, MØ05a, MBT^{+}03, ML04, MY06, MPdSM07, MNT95, MWY01, MVW09, MR97a, MMP03, MT02a, MdlBP05a, MdlBP05b, MRU12, Mar93, MBMS07a, MBMS07b, MOFM10, MTCOBH11, MBVTY11, MDY07, MJR04, Mau91, MPPJ^{+}05, MZCC09, May93, MST95, McA00, MBP^{+}11, McG08, MMM10, MFG12, MXZ05, MMO^{+}11a, MMO^{+}11b, MT08c, MCA^{+}07, MH04, MN98, MCSN10, MM04, MCP93, MS95, MMMD97, MdPA98, MSA99a, MSA99b, MP01a, MP01b, MA03b, MA03c, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, MGH09, MMK08]. June [Moh91, Moh97, MVA01, MMM05, MM97, Mor95, MD99, Mor06, MO05b, MR11, MSX10, MB98, Mue06, MCdS^{+}08, MdSC09, MV06, MMM11, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, MV07, Mur99, MO92b, Nag90, Nag95, NDFMS97, NJ10, Ng79, NGHJ07, NBGA08, NS00, Nie03, NCJK08, NPS06, Nob12, Nol81, NS90, Nyb98, ONC^{+}11, O’S03b, OFR09, Ode04, ORS89a, ORS89b, OKHH^{+}08, OYM07, OA07, OP97, OFK05, ORB^{+}09, Ong11, OKdQ09, OA93, OKBMB10, OPKR05, PR08, PRV11, PHR09, Pai09, PP01, PDH99, PDH01, PDHS03, PAD^{+}08a, PDML11, Pal00b, PZE^{+}08, PP12a, Par85, Par98, PCA^{+}09, Păs09, PF05, PSE10, PH10, PCPS03, PT98, Per10b, PS04b, PVTG07, Pet92, PS96b, PK12, PH06, PK00, PHOK10, PVV10, PP11, PS98b, PBRB08, PISW05]. June [PZ07b, PR98a, Pre09, PWY08, PM09, PRR05, PS09c, PB92, QLXL07, RKS89, RMDS10, RFBW12, RLL04, RM02, RK93, RM96, RS99b, RV10, RR01, RTA95, RA00, Rei88, Rei89, REF^{+}06, Rig08, Roc07, RK02, RKW04, RBC93, RV11, RCS10, RHD08, RJS^{+}08, RS03, RFBD09, RB00, RSTZ04, RTZŻ06, RTZZ08, RST^{+}10b, RST^{+}10a, SS07a, SP09, SNČ^{+}12a, SM11a, SS11a, SHN07, SHJ09, SKMR94, SZR^{+}08, SPCG07, SKR11, SHM^{+}10, SINK08, Sav06, SPH07, SH08, Sch03a, Sch11, SLZH07, Sel86, SO09, SMP07, She03, SF08b, Sif90, Sir12, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, Smi06, SCZ10, SMF^{+}11, SMMG01, SS04, SZ05, SRT06, SD10, Str96, SM01, SHYEF08, SHUH08, SKN05, SKR^{+}10, TMLL09, TST10a, TST10b, TSCW11a, TSCW11b, TCHB12b, TCHB12a, TSJ12a, TSJ12b, TBKO10, TY11, Thi10, TS87]. June [TBJ02, TBG05, TN09, Tom92, TT06b, TG10, Tre09, TV87, TSKGB04, TAS05, UBM02, Vah09, Val94, Van89, VMvN12, VW05, Vau06, Vau08, Ver00a, Ver00b, VV07, Vit10, VSH11, Vos03, VWY97, VWY01, Wad87, WAS06, WR01, WLY05a, WLY05b, WLY05c, WYZ06a, WYZ^{+}06b, WB00, WSTI07, WNE99, WP10, WBW94, WR03, Win11, Win09, WZA10, WHS01, WANL08, XCd03, YBM87, YD98, Yal10, YCC^{+}08, YAKMJ11, YKL06, ZYRP10a, ZYRP10b, ZCCL10, ZLK10a, ZLK10b, ZS10, ZYB06, ZY10, ZHP12, ZYW08, dBNT87a, dBNT87b, dBdRR89, dBdRR90, dBdRR91, dB92, dBdRR93, dBdRR94, dMB05, dS11a, dPMA98, vEGW09, vHV08, vL89, vL94b, vO04, vdAB03, vdAHW03, vP93]. June/July [BBP93]. June1 [BJP12]. Juni [KF74]. Jürgen [BKT12]. JURISIN [NMM10, OBS12, OBM11]. JURIX [CPSA10]. Jyväskylä [AM97, BM98, ILR95, PB08b].

K4CARE [Ria08]. K4HelP [Ria09]. Kōbe [FvdHJT10]. Kagawa [MF08]. Kailua [WD10a, WD10b]. Kailua-Kona [WD10a, WD10b]. Kaiserslautern [AFM01, ADB^{+}05, BLKP08, BKS01, BBR90, Bra75, DBB^{+}08a, GR06b, GNR94, Hsi95, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, Neh88, Nie93, RB91, RB00, SSW92, Sch12, SB07, Sti90, WAR94b, Wet92]. Kalamata [GMV05]. Kalamazoo [SdDK91]. Kamakura [IZ99]. Kananaskis [KH02]. Kanazawa [FM96, GMO09, IK12, KS05a, NY93, SHN07, vdHIP11]. Kanpur [AS02a, DBC^{+}10]. Kansai [BI04, ZJS^{+}02]. Kansas [Mel86]. Kaohsiung [PCN10a, PCN10b, PCN10c, PCN12a, PCN12b, PCN12c]. Kaprun [Con00]. Karlovy [MMP05]. Karlskrona [JC12]. Karlsruhe [BS09, BPR08, BWAM09, BCRB^{+}11, BB89, BG11b, BBR88, CC93, CL96, CL08a, CSR08, CL08b, Deu81, DBHS10, DJS91, Gel99, HM08, KW07a, KF74, Nic75, Ott87, PT92, Sch93a, SUW02, SSD^{+}08, WHS01]. Karpacz [MT04a]. Kassel [GKMV11, SNZ08, SDM05]. Kathmandu [MAD^{+}04, NRR12, SU97]. Kaufman [Gol11b]. Kazakhstan [VV03]. Kazan [AM10, BBL87, Mal11]. Kazimierz [PT95, Tar91]. KBCS [RCA90]. KCSD [NMM10]. KDD [BFMS08, BFJM09, GVO^{+}10, ZSM^{+}09, ZSD03]. KDECB [TWSN07]. KDID [BB06, DS07, GS05]. KDLL [BHH^{+}06]. KDMCD [ZSD03]. KDO [ABG^{+}06]. KDXD [NZ06]. Keele [GW93]. Kelowna [GJ09]. KELSI [ABD04]. Kemer [ADK04]. Kemer-Antalya [ADK04]. Kendal [KFLS07]. Kenting [JS11b, Yan11b]. Kerkrade [KL07b, NSM00]. Kernel [SW03a, SW03a]. KES [AHJ07a, AHJ07b, AHJ07c, GHJ06a, GHJ06b, GHJ06c, HNH^{+}09, JNHJ10a, JNHJ10b, JKN^{+}12, KHJ05d, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, LHJ08a, LHJ08b, LHJ08c, NHJ04a, NHJ04b, NGHJ07, NJHJ08, ONC^{+}11, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, VRHJ09a, VRHJ09b, NHJ04c, PHJ03b, PHJ03a]. KES-AMSTA [HNH^{+}09, JNHJ10a, JNHJ10b, JKN^{+}12, NGHJ07, NJHJ08, ONC^{+}11]. KeY [BHS07, AL06, AKTT12, BDZ04, CL11, CFGN11, Cra08, Des02, FBM12, JT09, Jue04, KGL04, LSF07, MP10b, MMK08, NP10, OW07, PNPP12, YDKM06, CZ05, IZ98, IZ99, IZ00, Kim01a, NP02, Vau05, WM04, BFS92]. Keynes [GHK^{+}10]. KGC [BM03a, GLM97]. KGC’93 [GLM93]. Kgs [VTK^{+}12]. Kharagpur [CRD08, PKKG06, RS97a]. KI [BFP04, DBB^{+}08a, FKS07, Fur05, HBE07, BBE01, BE11, BFP04, BHN97, BCC99, DBB^{+}08a, DBHS10, FKS07, Fur05, GK12, GH96, GKN03, HBE07, HG98, JKL02, MHA09, NDF94, RBW95]. KI-94 [NDF94]. KI-95 [RBW95]. KI-96 [GH96]. KI-97 [BHN97]. KI-98 [HG98]. KI-99 [BCC99]. Kibbutz [EKM06, ER12]. KICSS [TKSN11]. Kiel [BHL90, KK97, LS12b, SF74, SK97, SDP97, SZ00]. Kiev [Pfe94]. kinds [Wag94]. Kingdom [BGM94, Bow94, CC95, DCP95, EHRT08, GK95, Gol94, JBJ94, Lom97, Lou94, ML95, MK96, PRJ95, RT95, SOS96, Sha86, Sma97, Wat95]. Kingston [Akl95, ACD^{+}07, BK02, DOSY05, GP08b, HA00, MCSN10, PT06a, TM99]. Kinloch [CM00]. Kinross [Bak95]. Kinsale [Cag02, Fos02, Ros03]. Kirchberg [GAR04, Gue05, GRR05, GAR03]. Kish [AS10]. Kitakyushu [IDMY08a, IDMY08b, LMA08, TNY07]. Kitakyushu-City [LMA08]. Klagenfurt [ALS^{+}05, Bos96, BS03b, Bou11, BFOY04, DKM^{+}05, EM03, KBHSW06, KFMM06, KBH03, LHH10, Mit05, SMH^{+}12]. Klagenfurt/Velden [EM03]. Kleene [BMS04a, BMS08b, BJM09, MWD06, Sch06]. Kleene-Algebraic [BMS04a]. Klein [Ley01]. Klop [MvOvRdV05]. Kloster [ANR99, ADMH04, ADM^{+}06, ADM^{+}08, GS88, Mol95, RABR03, SSU12, SH11]. Klosterneuburg [CH10]. KMGov [Wim04, Wim03]. know [Sin94]. know-how [Sin94]. Knowledge [AS11a, APHA11, ABD03, ABT04, BBGS10a, BBGS10b, BBGS10c, BS11b, Bar02, Ber12, BS10, BF06, BMYZ95, BEGP04, BJ01c, BGMST09a, BGMST09b, CLY02, CRW^{+}12, CD11, CD12, DGM08a, DGM08b, DSZ04, DT07, ES02, EH08, EMT02b, FDC12a, FDC12b, FSS06a, GTWC09, GPB02, GJS08, GHMV11a, GHMV11b, GHMV11c, HKW09, HKW^{+}10, HKW11a, HKW^{+}11b, HKW12a, HKW^{+}12b, HKI08, HCL05, HhKRT06, HCS11a, HCS11b, HKH10, JTB^{+}05, KMT00, KMW03, KMW04, KR02a, KR04a, KKdM^{+}07a, LGTB03, LMA09, LNW^{+}12, LFK^{+}02, LP10b, LS12b, LCDT08, LDC^{+}09, MS10d, MLK04, MdPA98, MÁ05a, MÁ07b, MNLM06, MSS04, NZS^{+}06, NSS^{+}07, NHJ04c, NT06, NKLC06, Paz03, PMT09, PMT10, Per02a, PPH07, Pup99, RK06, Ria08, RK09, RK12, RB00, ST08, ST11, STT04, SHW09, SZC^{+}03, SEN07, SEN08, Sta99]. Knowledge [SS06c, SPS09, TCHB12b, TCHB12a, THL11, TDG91, TKCH09, TT06a, TO08, UWB^{+}06, Urs02, VY93, WJ05, WJS^{+}06, WPSY06, WLGB^{+}08, WZH^{+}07, WSTI08, WLP^{+}09, XCR11, Yam08, YLW^{+}07, YRWS11, YGL^{+}10, ZYRP10a, ZYRP10b, ZSM^{+}09, ZLY07b, ZKZ00, dMLG03, dSOSR04, dSOSR06, tTVH^{+}12, vEDA04, ADB^{+}05, Aus93, Bar03, BMYZ91, CWL01, DC00, DPP98, DH06, EEM96, FNY93, HNR05, JL95, KWA01, KWA02, KL95, Koh06a, KZ97, LN94, LCS91, Mat91, MMS93, MT99b, PB92, RB91, ST00, SSW92, SHGB05, SOS96, SXC04, SSv94, SDM05, TT05, Wag94, WP90, Wet92, Wha03, WKK98, X^{+}12a, YBM87, ZZ99, ZQ98, ŻR99, dRS01, ABD04, AHJ07a, AHJ07b, AHJ07c, Aus93, BW10b, BB06, BEGP04, BHH^{+}06, CP10c, DC00, DS07, FS99a, FHW98, FSS06a, GVO^{+}10, GHJ06a, GHJ06b, GHJ06c, GE08, GS05]. Knowledge [JTB^{+}05, KR10a, KJ09, KHJ05d, KHJ05a, KHJ05b, KHJ05c, KKdM^{+}07a, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, LLW06, LGTB03, LHJ08a, LHJ08b, LHJ08c, MP94, MB09, Mor89b, NZ06, NHJ04a, NHJ04b, PHJ03b, PHJ03a, PST11, PB97, RCA90, Ria08, Ria09, RtTMP10, RtTMP11, RtTM12, RK09, RK12, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, TLC00, TK05, TKSN11, TWSN07, VRHJ09a, VRHJ09b, Wim03, Wim04, WPZA11, XL11, ZS07, tTVH^{+}12, Her94, PZ92]. Knowledge-Based [BMYZ91, BMYZ95, HKH10, MdPA98, NHJ04c, VY93, LMA09, Pup99, Aus93, JL95, Wag94, WP90, YBM87, AHJ07a, AHJ07b, AHJ07c, GHJ06a, GHJ06b, GHJ06c, KHJ05a, KHJ05b, KHJ05c, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, LHJ08a, LHJ08b, LHJ08c, NHJ04a, NHJ04b, PHJ03b, PHJ03a, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, VRHJ09a, VRHJ09b, KHJ05d]. Knowledge-Centered [HKW09, HKW^{+}10, HKW11a, HKW^{+}11b, HKW12a, HKW^{+}12b]. Knowledge-Driven [PST11]. Kobe [CDP02, EKS10, GMV09a, GHH^{+}06a, ST12a]. Koblenz [Bec05, DHH^{+}08, EKL^{+}05, Fur05]. Kolimpari [PR11]. Kolkata [AKG06, Asa06, BA06, BL06a, DD03, DSDS04, DU09, DLW12, GN12a, GDP07, JM05, JM11a, KPSW10, KMMP12, PBB05, PG09, RCJ^{+}08]. Köln [Sch74]. Komárno [HPZ10]. Kombinatorik [BHR84]. Komplexität [SS76]. Kona [WD10a, WD10b]. Kong [BDDM12, BCF^{+}02, yCW96, CYGL04, CNX05, CSJD06, CT06, CTS10, CP06b, CWL01, CY12, Chi95, CP98b, CP98c, CBS^{+}11, CCAS12, DW07a, DY06b, FT04a, FCLL02, FKW08, FHW08, GPL99, HL00, cHPkC^{+}07, IMY^{+}07, IS98, IAL^{+}07, KWCW06a, KWCW06b, KWCW06c, KYT01, KZ10a, KFfKL11, LLCL05, Leo00, LwCM00, LPC^{+}11, LMA09, LCB12, LYhL^{+}01, LCY03, LL06, LM08b, Luo11, Ng93, RWA06, SM11a, TLF01, TWYhL01, XYM^{+}07, XYZU11, YKF^{+}06, YKL06, YKU11a, YKU11b, ZJ04a, ZJ06, Zha07, ZS10, ZYW08]. Kongens [Kin08, NF07]. Konigswinter [EW91, SP07]. Konstanz [BGD^{+}05, BW00, BC11c, BE01, FM91, Fum97, Zan00]. KONT [WPZA11]. Korea [ACN^{+}07, AYBH^{+}06, ACP05, Bai05, BL10b, BBK^{+}11, BCDN07, CCK^{+}08, CY03, Cha02a, CCP10a, CCP10b, CPC07, Cho02a, Cho02b, Chu03, CSY09, CY11b, CI98, CM04a, Col07, CKL05, DWSK05, DKY06, Dat09, FKLY08, Gel04, GTM^{+}09b, GTM^{+}09c, HSSY05, HK05c, Ho12a, Ho12b, HTMC09, HI10, HYPY10a, HYPY10b, Hu09, HJL^{+}09a, HJL^{+}09b, JP06, JY12, Kah03, KG04, KCR04, KR10a, KM96, Kim01b, Kim01a, Kim02, Kim05b, Kim05a, KT06, KYL07, KGV^{+}09, KKH09, hKhLKŚ10, KSB11, hKAS^{+}11, Kim12, KDWW10, KBC05, LBP00, LB03, LL03, LhK03, Lee04, LLWL04, LYMM05, LL07, LKK^{+}07, LYY07, LCHS08, hLhKcFŚ09, LC09, LBM^{+}10, LH10, LW11, LHŚ11, LHKŚ12, gLPZ^{+}12a, gLPZ^{+}12b, LY12, LL04, LY05, MY06, MTU09, MBJ^{+}07, NR07, NJHJ08, NKJ11a, NKJ11b, OB07, PC05, PCA^{+}09, PZYS12, PK03, RL06, RN11, Sch11, SHX^{+}06, SJ06, Son00, SKY06, SKCK07, SHŚ^{+}07, Wha03, Won01, WK06]. Korea [YMHH10, YKF97, Yi06, YY09, YKM06, YYH^{+}12, ZO10, ZSY^{+}06]. Kornik [Kar77]. KORSO [BJ95]. Kos [PGA^{+}05]. Kosice [PRR94]. Kowalski [KS02a, KS02b]. Kozen [CS12b]. Kozubnik [KMS89]. KPP [WPZA11]. KR4HC [RtTMP10, RtTMP11, RtTM12]. Kraków [BvADS08c, Abr01, BPW06, CC97a, SSZ12, TT06a, BKW06, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BvADS08a, BvADS08b, FHFL06, SCKR10a, SCKR10b, WSGA06, ZCCL10]. Kralendijk [BDC12, Ker12]. KRAMAS [MB09]. Krämer [Hei12]. Krasnoyarsk [Mal05]. Krems [PMD92, Wim04]. Kreuth [BEM91]. Kristiansand [SJW^{+}12, TMS12]. Krumlov [Kuc̆02, KK07]. KS [GH12a, MCSY07]. KSEM [BW10b, KJ09, LLW06, XL11, ZS07]. KSinBIT [MTH06a, MTH06b]. Kuala [CWYC12, DV05, GG07a, GG07b, GG07c, HWG10, LM09, LS09, MZ99, SZC^{+}03, TCHB12b, TCHB12a, ZRP^{+}09]. Kuching [AIL12, CFL10, Kur07, RWSS12, RK12]. Kühlungsborn [UBM02]. Kunming [GSV05, HLI06a, HLI06b, HLI06c, SWJX12, Wan05b, WZH^{+}12, WYX08, ZYH03]. Kuopio [KTB09]. Kuressaare [JV06, Uus06]. Kurt [AAN09, BM03a, GLM93, GLM97]. Kusadasi [DGR06]. Kusatsu [CKOB11]. KY [GPPW08, Die03, MNT95]. Kyongju [KM96]. Kyoto [AT12, AM00, BFK^{+}07, Got83, GAY86, HM12b, IKO03, IGN05, IFV07, IT03, IKKU08, IT08, LO98, NcC99, Oka00, OA07, Pol99b, SNE08, SHRM06, TPB79, TvdBI02, TFK11, TLC00, WSS04, Wei97, YM01, YSR^{+}06]. KyotoCGGT [IKKU08].

L [TDK^{+}10, BBB^{+}85, RS74, Woo80]. Lac [CFG96, JP11a]. Lac-Beauport [JP11a]. LACL [BD12, BSBM05, PP11, DMR01, LLP99, Ret97]. LADC [BBR07, MSAdAS05, dLWC03]. LADS [DSLT08, DSLT10, DSHL11]. Lafayette [EV01, ELM05, EdS08]. Lagos [Bra07]. Laguna [HM96, KS07b, LGM11]. Lake [BBKP06, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, CW00, FKM08, HK07, LLS00, LMT00, Mil01, SS07a]. Lakes [SdDK91]. LALC [Moo01]. Lam [BG93b]. LAMAS [TtHVS06]. Lambda [Boh75, Cur09, Ong11, Roc07, Abr01, BG93b, DH97, DCP95, Gir99, Hof03b, Urz05]. Lamia [MPV12]. Lancaster [Mez11, SBC^{+}94, She01]. Landscape [MP06]. Langhorne [FGH98]. Language [AFP^{+}10, ABdJS10, AKR^{+}12, ABJ02, BSMM04, Bar08b, BBB^{+}85, BdQ11, BEP04, BEJvR11, BCG^{+}11, BLSS11, BM03b, BGL09, BIMW12, CVTP12, CD00c, DdQ10, DN09, EJM10, EKF12, EKS00, FGK^{+}11, FMPT10, Fuc10, GLSY10, GLV09, GR06b, HKL07, HdQ08, HRM^{+}10, HMMW10, HMCL06, ID06, IM10, IK12, IT08, JHC02, KKKS11, KSS08, KLM^{+}07, Kon04, KFMM06, KT10, KF12, LS12a, LdQ07, LMA09, LRH10, MBS11, MBT^{+}03, MM09a, ML11, MM04, MMM05, MR12b, MMM11, NR08, OdQ12, OKdQ09, PPCS12, PBK^{+}10, Paz03, Pet01, PBGK03, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, RM02, RF12, SGPP06, SA12, SWB03, TDB^{+}06, TdLdOQ08, Uni81, Vää99, VU09, Vet11, VMBMN04, V^{+}06, YI12, YJKK06, YJKSW10, tCZ07, vdBGG10, AZ96, Bar97, BM99b, BJ78, Bla88, Bol78, BKM01, BLW82, Bre97]. language [BW92, BP93, Bus85, Bus88, CCD04, CSV05, Chr00b, Dal92, DWSK05, DR05, DM90, Don76, EK77, EF03, Fil91, FR99b, Fra96, GK01, Goo83, Hau86, HP94, Hir81, HKB93, Hou90, IT03, KSW90, KRS02, Luc87, MML97, Mai88, MW91, Mar80, Mit98, PBGK02, PGBK04, RS96, She85, SG96, Stu90, STLK05, TD97, TDBP01, TJHS94, UA83, WF98, WS02, WSL93, YI91, BS76, BEJvR11, BBPP09, CW02, DIMV09, DFMV10, DIMV11, DMV12, FMPT10, MVOF08, Mos04, Tob80, Pet01]. Languages [ADG^{+}08b, AHS11a, AHS11b, AST05, AMSM^{+}09b, ABDW10, AB04, AS07b, BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, BdBC01, Bar11, BGS^{+}03, BI73, BS03b, BPSW06a, BPSW06b, BGH06b, CP10b, CSCR08, Cas09a, Chi04, CHZ08, CW06, CA10, CM00, COB^{+}08, CMPW12a, CMPW12b, DR11, DFM04, De 07b, DO89b, DKLS04, Dig04, Dro08, EOSW07, FV09, FKBA12, GG09, GS09b, GH12a, Gor10, HM11b, Han07, Hen06, HZB07, Hu09, HW08b, Hug03, Jay04, dCJB12, Kob06, KR02c, LS04a, LZ08, LOST04, LOTY05, LMA09, LS06a, MSWZ06, MS11a, MV07, NWHR06, NJ09, NJA^{+}11, ORS89a, ORS89b, Oho03, OSV11, PA01, PRH10a, PRH10b, PC00, PR00, Pri03b, Ram08, RL11b, RZ12, SST00, SSVW12, Sar03, SF74, Sch04b, SC11, SW03b, SS09b, Sei12, Ses06, Sha07, Sir12, SWB03, Tah09, Ued10, WCK11]. Languages [WTM^{+}02, WJ95, Yan11b, dBNT87a, dBNT87b, AvLR91a, AvLR91b, AS94, AGK^{+}10, ADG^{+}08a, AMSM^{+}09a, AT02c, ACJT07, AM02, AB78, ADCR89, BSMM04, BLPW03, BCB99, BL92, BJ84, BK05b, BRW93, Bra75, Bra85, BW05b, BJ95, CIM^{+}05, CL01, CNY95, CH96, CW99, CHD98, CH98, CCR86, DW03a, Dam97, DK89, DGMS97, Deg03, DLM89, Dia83, ES92, EK81b, FG95, GL97, GG02a, GHK97, GK01, GHMT05, Gup99, Gut94, GW00, Hal94, HI93, HDC99, Han98, HP95, HS95a, HC05, Hug91, IH90, IHQ96, JP05a, JSdSA05, JL00, JM78, Jou85a, Kah87, Klu97, KC00, Kot86, KS96, Kui92, LSW98, Le 02, LS88, LKC93, Loe74, Mau79, MT02b, MM96, MK01, MAA91, Mos97, MWJ97, MSR99, NS82, Nie96, OSvL01, Ott87, Par84, Pat90, PS97b, PA03a, Ram01, Sag05]. languages [SS77, San01, San94, SRW98, Smo00, Swi99, TMP04, WMR00, WF77, WMT96, Yi05, ZL98, dBdRR91, vWN99, vLdB80, AGP08, ACW07, Ama08b, ABRS06, Ban92, Ban93, Ban94, BKT12, CF00a, CPC^{+}99, CMCS11, DSLT08, DSLT10, DSHL11, DM01a, Die03, Dwa00, ELM05, GPCL10, GKKW09, HSB^{+}96, Li98, MMG^{+}01, MB98, O’H98, Pin95, PT06b, Rau04a, Seh97, Bee86, Eng74, Hab92, Kli85]. Lanka [MVCM10]. Lans [War94a]. LANSEC [BB89]. Lanzarote [MMMD97]. Lanzhou [HLCZ11, PF07a, ZCGH11]. L’Aquila [Gir99, AP83, GT03]. Large [AMZ09, DKK08, DFP^{+}12, GLZ^{+}03, HKW09, HKW^{+}10, HKW11a, HKW^{+}11b, HKW12a, HKW^{+}12b, LWZ09, Mei08, Pol01, TO08, Buc90, Hil78, Kåg98, KvG98, ZWS81, CG12, Col84, LMWY04, LMW06, LMW08, LMW10, LMW12, MWY01, SvdMOP06, TO08, ZH00]. Large-Scale [AMZ09, DKK08, DFP^{+}12, GLZ^{+}03, HKW09, HKW^{+}10, HKW11a, HKW^{+}11b, HKW12a, HKW^{+}12b, Pol01, TO08, KvG98, CG12, LMWY04, LMW08, LMW10, LMW12, TO08, ZH00, LMW06, MWY01]. LASER [MN12a, MN12b, Mül10]. Lassez [Mah04]. LATA [DIMV09, DFMV10, DIMV11, DMV12, MVOF08]. Latent [SGG06, TCYZ12, VZM^{+}10]. Latin [AB10, BYGP95, BBR07, CHK06, FC04, FB12, GPV00, HN12, LBNF08, LO10, LM98c, MSAdAS05, Sim92, dLWC03, BYGP95, CHK06, FC04, FB12, GPV00, LBNF08, LO10, Sim92, Raj02]. Latin-American [BBR07, MSAdAS05, dLWC03]. Latin’98 [LM98c]. LATINCRYPT [AB10, HN12]. Lattices [Ekl04, YNB08, Sil01]. Latvia [AF06, BDGM05, BGHS05, Fre01, GMV09b, GKMN10, MA07c, PS04b]. Lauderdale [DL04b, LN03a, MNZ09, OPS11]. Laurel [HF97]. Lausanne [CG09, DH07b, ES10, FPFR05, FMN99, GGHN97, IMW04, JNS05b, KTC^{+}05, LV05, Nyb08, RG05, Sch93b, SMPU98, SF96, TMG10, VDÉ^{+}12b, VDÉ^{+}12a, WCJS10, WD00]. Law [CSCR08, FMPT10, BCBG05]. Lawrence [GH12a, MCSY07]. Laxenburg [BMW10]. Layer [Har11b]. Layered [Pup10, Mul96]. Lazne [Bec75, SPK01]. LBJ [CW00]. LCC [Lei95]. LCF [GMW79]. LCPC [AGP08, ACW07, Ama08b, ABRS06, CMCS11, ELM05, GPCL10, MMG^{+}01, PT06b, CPC^{+}99, Die03, HSB^{+}96, Li98, Rau04a]. LCPC’96 [Seh97]. LCPC’99 [CF00a]. LCR [Dwa00]. LCR’98 [O’H98]. LCTES [DM01a, MB98]. LDAIS [Wei97]. Le-Mont-Saint-Michel [HR95]. Learning [AKZ01, AJS00, AM05, BLS06, BBGS10a, BBGS10b, BBGS10c, BDCM04, BEG04, BvLR04, BG07b, BOH^{+}12, BGMST09a, BGMST09b, BSG03, BSPB07, CBNR02, CKK^{+}09, CW07, CFK^{+}12, CYBF06, CY09, CDS09, DGM08a, DGM08b, DBC^{+}10, DJW04, DS08, DGDM^{+}11, DKW07, EMT02a, FDC12a, FDC12b, FCLL02, FKW08, FNGW08, FGTZ08, FSS06b, FKLY08, FTC^{+}10, GCB^{+}05, GJT03, GLZZ09, GVB12, GLM^{+}08, GHMV11a, GHMV11b, GHMV11c, HM03, HDGK11a, HDGK11b, cHPkC^{+}07, HST07, HSVZ10, KS02d, KSUZ11, KGG^{+}11, KKdM^{+}07b, KFfKL11, LSW00, LGBT03, LSR03, LHH10, LwCM00, LWM^{+}08, LLLL08, LPC^{+}11, LZS^{+}08, LKZ^{+}08, LCY03, LLL06, LS06c, LSW^{+}10, LCY^{+}11, MM08, MH04, MP01a, NT06, PKS01, PAD^{+}06, PZE^{+}08, PI05, Per07b, Per09b, Per11b, Per12b, PBSB09, PRG09, PRG10, PRG11, PBRB08, PBS08, PLK^{+}12, QCDM06, RLKHL12, RBF06, SH12a, SW03a]. Learning [ST12b, STS04, SLKL09, SG01, SWSY11, TT10, TCLH10, TNGK08, Ull08, VDÉ^{+}12b, VDÉ^{+}12a, VKG12, WLC^{+}06, WFZL09, WYSS10, WPE05, WNL05, WKS^{+}10, YEY04, YLWY06, YAF^{+}02, YTC^{+}07, YWRS11, YCB12, ZNCF03, ZW09, AFM01, AJ94, BEM91, BD97a, BB05a, BR06, BdR94, Bra93, B^{+}12b, CHCM11, DF96, Dos93, FS99b, Fre03, GHM05, HRR93, HW01, Jan93, JL95, JSZ95, KW98, Kod91, LSW01, LSW02, LLCL05, LW95, LM97, LSL04, LP00, MMR07, Mau89, McK99, MS03c, MD96, MSS05b, NR98, NS90, PP99, Per01, PR03b, Ric98, RS96, SA96, SJT05, Tom92, Vel94a, Vit95, WY99, WS96, Wei97, WD00, YKF97, ZZP^{+}10, dRF01, vW97, AGT98, AM05, BBMB^{+}08, BBD^{+}10, BD98, BB10, BG07b, CKK^{+}09, Coe11, CDS09, CD00c, GMUW12, HS12b, KLT^{+}07, LSW00, LSW03, LS06c, MBW08]. Learning [Pau01, RB00, SW03a, STS04, Stü09, TtHVS06, AKK03, Fre03, KKA05, Mor89b]. Learning-Base [CW07]. Learning-Based [Pau01]. learning-ICWL [LSL04]. Least [Ahn04]. Lecomte [PQR11]. Lecture [BG12a, Lau93a]. Lectures [BvLR04, CEFZM05, GAB02, Mül10, MS03c, MSS05b, TBG05, ACP01, AGM05, AG07, ABG09b, ABDW10, AG11, Alt01, AAB^{+}07, ABW10, BG03a, BGHJ07, BBF^{+}06, BBM^{+}08, BDPS02, BH07, BDZ08, BPZ09, BI11, BCP12, BIK01, BG12a, BCSW10, BC08, BvLR04, BBPP09, BHK01, BJK^{+}05, Buc85, CT02, CG04, CSW06, Dam99a, De 87, DF09, DRR04, EM05, EFR08, FG01, GLW07, Gib12, GI03, Hor06, HPSZ08, HPZ10, IJ06, KEM07, KPS09, LABM^{+}12, MSY08, MPS98, MSS03b, MN12a, MN12b, PKS01, PdA^{+}11, PR98a, RR98b, RR98c, RG03, RV98, TFE^{+}09, WW07, WNL05, ZYW08, dBBGdR03, dBBGdR06, dBBGdR07, dBBGdR08, dBBM09, BCG02, BI03, BC04a, BB05b, CB06, CMT01, ÉEMS04, FG04a, GG98, JJ03, KSS02, MS03c, SHO99, Vää99, VU05, dBBG04, dBBGdR05, dRLP98]. Lednice [KBČ^{+}12]. Leeds [Fog94, WM03, vDKS11]. left [Hau86]. left-associative [Hau86]. Legacy [MFÁ^{+}09b]. Legal [BBGR03, CPSA10, FMPT10, PPCS12, BCBG05]. Leicester [EHRT08]. Leiden [Bos00a, CR01b, HPRS06, dBBGdR03, dBBG04, dBBGdR05]. Leipzig [BHU03, BLVV07, Dav89, DW10, DHSH96, Hof97, MP06, MSG07, MG08, PpWR96, PP99, Per01, PR03b, Per05, PI05, Per06, Per07a, PS07a, Per07b, Per08, PS08a, Per09a, Per09b, PMKG07, Rah04, WPR08]. Leisure [LHH10]. Lemesos [Pel07]. Lemessos [IFGH10]. LENLS [NMM10, OBS12, OBM11]. Lens [GG05b]. Leonard [MM09a]. Leonid [Gol11b]. Leopoldo [SVA05]. LerNet [BBPP09]. Lesná [ABH12, KU06]. less [FK09]. Leuven [AD11, BDE12, BPV04, BG08, BW92, CS01b, DR02a, De 10a, DIP12, Dun09, Fuc98, GANACBdCdV12, IS08, MRZ09, PVV05, PNPP12, PGV93, Pre95, PR98a, Pro11, PS12c]. Level [Mah04, PHSZ06, SHM^{+}10, Sch02, Sha04, BL92, BLW82, PFCS05, Coi99, Mue01, Sch91]. Leveraging [MS10b, MS10c, RDK09, MSG07, MSS12, MS06a, MS10b, MS10c, MS12b]. Levin [Gol11b]. Lexical [PB92]. Lexicalized [Kuh10]. lexicon [Ste95]. lexicons [Ste12]. Lexington [GPPW08, MNT95]. LFCS [AN07, AN09, AN97, NM94]. Libraries [ABK^{+}09, BCG^{+}11, BB00a, BMC08, CCM^{+}04, CC12, CKH10, CDCJL08, GCSR07, GTVC06, GBMS11, KS04b, KFM07, LJR^{+}10, LFK^{+}02, NS98, SZC^{+}03, SHRM06, TDB^{+}06, TBC07, XCR11, ZBRL12, AV99, AYB95, Ada96, Bar97, BBEdK98, CS01a, FNPW05, HL04, PT97, RCT05, TD97, TDBP01, Bar08b, BC02]. Library [CR05b, TAS05]. LID [PZ96]. Lidor [Gol11b]. Liege [Wol95, FSP09]. Life [ARC^{+}07, BCBF08, BCD^{+}03, BGF06, BR12b, BHH^{+}06, CA06, CBT07, DT07, KKS11a, KKS11b, KK11c, KRH09, LK10, LHH10, LNE06, LFIM07, LFJI10a, LFJI10b, LJS^{+}10, LR05b, OYM07, PMH09, Rah04, RAW07, ABD04, BGD^{+}05, CFB^{+}05, FMN99, Gom01, KS01, KS05a, Mor95, PSD^{+}07, PSM^{+}07, LLIH07, LFJI10a, LFJI10b, LJS^{+}10]. Lifschitz [ELLP12]. Ligure [Arm02, DPP95, GT04, MR97b, San92]. Like [SKS^{+}09]. Lille [BD97b, KU09, RT00, TLC06]. Lillehammer [KDS01]. LILOG [HR91, HR91]. LILOG-project [HR91]. Lima [BO00a]. Limassol [APPE09a, APPE09b, IFL^{+}12]. Limburg [HKP11]. Limerick [BVO10, BH95, BS11f, CW01, HN06, OSR^{+}02, PP01]. Limoges [Gab13]. Lindau [KK98]. line [Pfl02]. Linear [CHG^{+}07, FZJE^{+}05, FZJE^{+}06, Koj91, ZWS81, dBdRR89]. Lines [BL10b, Cha02b, Nor04, eOP05]. Linguistic [LSR03, VSTN06, GM00, LCS91]. Linguistics [BD12, Gel04, Gel06, Gel07, Gel08, Gel09, Gel10, Gel11b, Gel11a, Gel12a, Gel12b, HKS09, Jha10, KH09, PP11, Vet11, BSBM05, DMR01, Gel01, Gel02, Gel03, Gel05, GP89, HR91, LLP99, Moo01, Ret97]. link [LMŠT01]. Linking [BS10]. Linkoping [Gyi96, AAR05, BR12a, Fri93, Kir96, Nie96]. Linz [BKW09a, Buc85, BV94, BC04b, Cav85, Coi96, DSB10, DKN^{+}11, DB09, FM04, FHLP09, GN05, GM05, HT09, HM09, KS93, KKDV02, MT09a, MSZ09, MKZ02, MKZK06, MKZK08, MKPZ12a, MKPZ12b, Mos97, PMT09, Pic86, Smo97, VT97, WSJT09, Win11]. Linz/Vienna [FM04]. LIOME [Wei97]. LION [BB10, Coe11, HS12b, MBW08, Stü09]. Lipari [BC08]. Liptovský [VBCBS05]. Lisboa [DD97, FV09, Kos98, LLPH09]. Lisbon [AM03, AL04, ARC^{+}07, AP11, BSMM04, BBF^{+}06, Ber10, BB00a, CIM^{+}05, CV07, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CCdS03, CP10c, Coe98, CM05b, DH01, DS98, DGWB09, DLGJ10, Est98, Fia99, FCD^{+}04, GPAD08, Gen09, Gue99, Han98, JSdSA05, KK11e, LO02, LT05, MOP11, MSS07, MM89, MD99, Niv98, Oli00, PPP07, Pet01, RFZ03, RtTMP11, SO09, Ste98, SM09, Wel01]. lisp [IH90]. List [Gur04]. Listening [ECV^{+}10]. Literacy [Mit05]. Literature [BS10, BHH^{+}06]. Lithuania [CE01, Mit06, VMB^{+}12]. LITP [Cho88, CCR86, GP89, Gue90, Pin89]. Live [Mul93]. Liverpool [AD98, FvdHKL06, JSO11, KW07c, MS10e, dBM10a, dBM10b]. Living [BHV11, ORB^{+}09, PLMV12]. Livingston [MM01]. LIX [Nie09]. Ljubliana [SL99b]. Ljubljana [BSTL07, BG12a, DLŠ11a, DLŠ11b, EF12, RT12, TW12, dGN12]. LKR [TO08]. LLLL [NMM10]. load [BGW93]. LOCA [HKS06, SLP05, CQSS09, HSS07]. Local [Ayk07a, Nar01a, SBH07, SBH09, Hof03a, Wal99b, BB89, MSH85, MBS05b, Nar01b]. Local-Search-Based [Nar01a].

Localization [FK09]. located [Jon00].

Location

[BJSW09, CQSS09, HKS06, HSS07, SLP05]. Location- [HKS06, HSS07, SLP05]. Location-Based [BJSW09]. Lochem [KC00]. Locomotion [Tu99]. Łodz [SKN05]. LOFT [BLvdH10]. Lofthus [Hel94]. LOG [MRY^{+}11, KMSS02]. Logic [ÅBE12, AFR^{+}04, ABGL09, ABFR07, BM03a, BEM91, BPV04, BLSS11, BG12a, BvDMST11, BDR10, BPT06, BRST08, BKV10, BHR84, BBR88, BBR89, BBR90, BAL88, BGL09, Bra02, BJ01c, BL04b, CKS04, CCdS03, CFR92, CLS07b, Dau03, DV10, DFKM08, De 10a, DPP09, DFN10, DLGJ10, DP10b, DHMM96, DH07b, EKLM87, Eng81, ELLP12, Esc10, Ési06, FRS08, FPFR05, FPP11, FST09, FvdTDG12, FL11, FT94, FM91, GH08, GMP06, GM06a, GS10, GK09, GH93, HW06a, HG00, HCF06, HY03, IM10, IST07, KS02a, KS02b, KS04a, KM08a, KP05, Kuc11, LO11, LS12a, LT05, LTÅ^{+}11, LN04b, MT04a, Mar11c, MLM90, MMP07, MCA^{+}07, MR12b, MOTN07, MTNL12, O’S03b, OFR09, Ölv10, Ong05, SS08a, ST12a, SHGB05, SHW09, Sha86, TJHS94, TT06b, Vää99, Vor92b, ŽL08, tCZ07, vdMvdT08]. logic [AP96, BC89, BBP89, BG91b, BE87, BMP90, Bor91, Bor92, Bor93, BGM94, BW01, BJS93, BW92, BP93, Bun96, Bur91, CS00b, CJE82, CF00b, DLM89, DM90, DJS91, DPP95, DPP97, DL04b, DER94, Dur12, Dyc94, DHSH96, DF99, EHSH92, FR99a, Fre98, Fri01, Fur95, FU96, GO94, GZR02, Gol82, GMW79, GLM93, GLM97, GGS99, GLW88, Had94, HR96, HHM97, HS94, HG96, HK98a, Hee94, HP94, Hol89, HRA02, Jac91, JS94, KW90, KL92, KS93, KU01, LM93, LD97, LA94, Llo00, LCG97, LIH99, MW91, MP94, MS03b, MP96, MC94, MS99b, ML95, Mug97, Mul93, MRS97, NT98, NCW97, PT95, Pag98, Pet92, Plu90, Ral94, RM97a, RS99a, RS01, SS89, SH91, SWAF95, SBG04, ST87, Stu90, Wag94, Wan93, vB97]. Logic [ABdJS10, AKR^{+}12, Alp11, ACÇ^{+}12, BV05a, BV02, BS11b, BS11c, BGLT05, BBS07, BDL08, BdQ11, BLSS11, BV12, BLvdH10, BGL09, Bos00b, Bru04, CVV08, CV10, CDE^{+}07, Cod01, CFR92, CS12b, CD00c, DN07, DdQ10, De 10b, DF11, DL04a, DV07b, DFN97, DPP98, EFT01, ELS09, Eta05, ET06, FV10, Fle99, FT94, Fuc98, FTF88, FTF89, FTF91, GG05a, Gal97a, GVM99, GLP99, Han09, HHP09, HV06, Hil06, HW09, HdQ08, Kai11, Kin08, Lau01, Lei95, LdQ07, Len04, Leu03, LN03a, MNT95, MT89, NV01b, NPW02, OdQ12, OKdQ09, Pal03, PV00, PZ96, Pet02, Pfe94, PQR11, Pro96, Pue07, RS09a, Stu02, SV05, VV03, Vid12, Vor92a, Vor92b, Vor93, Wad86, Wad87, dlBP08, tCZ07, vDLJ11, Bor87, Dau03, Hai82, Har79, Koy92]. Logic-Based [BL04b, ELLP12, Alp11, De 10b, Han09, Kin08, Pue07, Vid12, Bos00b, Fle99]. Logical [CDE^{+}07, GM00, MT89, Yao82, AN97, AN07, AN09, BD12, BBLT06, BSBM05, DMR01, LLP99, Moo01, NT92, NM94, PP11, Ret97]. Logics [ACG^{+}07, BW03b, BNUW09, CS00a, CMT02, GTW02, MMT08, SB07, AH00, BDH^{+}99, BJ01b, CMT01, FO95, GHK97, GN98, GF97, HS95a, HM05b, Jou94, KS96, PW92b, Sch97a, dBdRR89, vGH96, APO96, AL04, DdCF98, FvdHKL06, Fle02, HLW08, JN10, Koz82, KC84, Küs01, MPP94, MB96, OAdGBP00, Par85, PW92a, Sal83, dCHM12, vE91]. LOGIDATA [Atz93]. Logik [BHR84]. Logistics [BHJ^{+}11, HSSV12]. LOGLAN [KSW90]. Lökeberg [CDNS00]. LOMAPS [Dam97]. Lomnica [Gru77]. lon [ETG02]. London [ARS03, AJS05, Atl08, BBMB^{+}08, BMP00, BE08, BBBB08, Bra09a, CMSM^{+}10, CP06a, CH01, CF00b, DJAP07, Fie02, GLSY10, GY04, GS08, HDC99, HAB02a, HS05a, HB97, Hil06, HP99c, IKR00, IS11a, JN12, KMT00, KSS08, LMR^{+}06, Lau01, LSE02, LCB97, Llo00, LCG97, MZ07a, MR11, OPP04, RW11, RWY97, RT95, Sha86, Sma97, TG04, TT06b, VZ99, WS02, WY98, WCBD08, YHR^{+}09a, YHR^{+}09b, YP01, vH83, vdBV01]. Long [HC05]. Longitudinal [DFGN12]. Looking [GCSR07]. Loop [SHM^{+}10]. LOPSTER [Lau01]. LOPSTR [Alp11, De 10b, Eta05, Han09, Hil06, Kin08, Pue07, Vid12, Bru04, FT94, Leu03, Pet02]. LOPSTR’95 [Pro96]. LOPSTR’96 [Gal97a]. LOPSTR’97 [Fuc98]. LOPSTR’98 [Fle99]. LOPSTR’99 [Bos00b]. LORI [HHP09, vDLJ11]. Loughborough [CHA03b, MBS05a]. Louis [EM08, HCS^{+}05, IJ06, Mah04]. Louisiana [LPA00, Mai88, Mai90]. Loutraki [Mak86]. Louvain [UPB07, Qui98, QS00]. Louvain-la-neuve [UPB07, Qui98, QS00]. Low [LJS08]. Lozenetz [MVW09]. LPAR [BV05a, BV12, CVV08, CV10, DV07b, FV10, HV06, SV05, BV02, NV01b, PV00, Pfe94, VV03, Vor92a, Vor93]. LPAR-16 [CV10]. LPAR-17 [FV10]. LPAR-18 [BV12]. LPAR’99 [GVM99]. LPKR’97 [DPP98]. LPNMR [BGLT05, BBS07, DF11, DFN97, ELS09, EFT01, GLP99, LN03a, MNT95]. LPSS [CFR92]. LSGRID [KS05a]. LSMAS [JSS08]. LSMS [LFIM07, LLIH07, LFJI10a, LFJI10b, LJS^{+}10]. LSO [AFM01, HM03, MH04]. LSSC [LMWY04, LMW08, LMW10, LMW12, MWY01, LMW06]. LTC [VU09, Vet11]. Lubeck [LdRV94, FDL10, GHMT05, LR05a, CBNR02, LSS02, RM97b]. Luca [Gol11b]. LUCAS [FKS86]. Lucca [MSB07]. Lucé [DGG^{+}10]. Lucerne [CK12a, HS07]. Lucia [DDS12, Dan12]. Ludics [LT11a]. Luigia [SS06e]. Luis [BL04a]. Luleå [OKBMB10]. Lumpur [CWYC12, DV05, GG07a, GG07b, GG07c, HWG10, LM09, LS09, MZ99, SZC^{+}03, TCHB12b, TCHB12a, ZRP^{+}09]. Lund [Joh03, LKC93, MSB12]. Lutherstadt [ZT04]. Luxembourg [Bir07, DP06, GAR03, GAR04, Gue05, GRR05, KHSM^{+}08, Luo09, RS09d, SNČ^{+}12a, vdMvdT08, GAR03]. Luxembourg-Kirchberg [GAR04, Gue05, GRR05, GAR03]. Luxor [BBT08]. LVA [TCYZ12, VZM^{+}10]. LVA/ICA [TCYZ12, VZM^{+}10]. Lymasol [PR00]. Lyngby [DW94, DMW96, DMW06, Jou11, Kin08, NF07, RR98a, VTK^{+}12, Waś96]. Lyon [ARSB09, AW12, BHRU09, BM94, BCRT92, BFMR96a, BFMR96b, B^{+}12b, CHAL04, CSTD08, CvRVSN12, DSHL11, DFPV11, FR95b, GLP12, GG09, GP09c, HRZK02, JP09b, Lac10, Lau00, Mag03, MMU96, NP09, RH01, ZKZ00].

M [Len04, NMP^{+}04]. M/E [NMP^{+}04]. M/E-Commerce [NMP^{+}04]. M2AS [SPC^{+}08]. MA [ACHH12, BBMB^{+}08, Con97, Dov05, FFRS05, GJRR08, GJRS12, Kil05, LKT08, LBDW09, LS10b, McG08, MMK^{+}05, NLS08, OSRS07, PZK07, RH98, SCHS07, AP04b, CL01, CKO01, CS02b, CM04b, DM01b, DKR02, FV01, FV03, FR01a, Fre96, Gae96, HI93, Hug91, JKRR04, JQ04, KK00, KP99, KP01b, KM00, Nao04, Nor04, Pic01, PC00, RV02, RPZ97, Sur02, WC98]. MAAMAW [CW94a, CM95, VP96]. MAAMAW’94 [PM96]. MAAMAW’97 [Bv97]. MAAMAW’99 [GB99]. Maastricht [CKTW12, DS08, KHL10, OFK05]. MABS [AT07, APN08, BGJ11, DS09a, DLT05, DP10a, HENR03, SA06, VSMS12, DM01b, SCG98, SBD03]. Macao [LH06, LR09]. Macau [CFLR12, JLW07b, WFZL09, ZTTY12]. MACE [BFvdM10, SGD09, vdMBD08]. Maceio [LVP04]. Machine [BdR94, BEG04, BvLR04, Bra93, CHD07, CMM^{+}09a, DGDM^{+}11, DKP00, DLO03, EMT02a, EBAH08, FGH98, FT04b, FSS06b, GCB^{+}05, GDP07, GVB12, HRR93, HDGK11a, HDGK11b, KKdM^{+}07b, KKDV02, KKD04, KMMP12, KMKP11, LK09a, LKD08, LW95, LGBT03, LP00, MM08, MTWD06, NR98, PRG09, PRG10, PRG11, QCDM06, Ric02, RWD09, VDÉ^{+}12b, VDÉ^{+}12a, Whi00, WNL05, YLWY06, ZJL06, ZW09, dRF01, vW97, AD98, Bod96, BDW97, CS87, CD01, DKD05, DLM99, MMR07, MS03c, Mor89b, NA97, PBB05, Wei97, BBGS10a, BBGS10b, BBGS10c, BB05a, BR06, BdR94, BEG04, Bra93, BGMST09a, BGMST09b, DGM08a, DGM08b, EMT02a, FGH98, FDC12a, FDC12b, FT04b, FSS06b, GCB^{+}05, GHMV11a, GHMV11b, GHMV11c, Kod91, KKdM^{+}07b, LW95, LGBT03, LP00, NR98, PKS01, PP99, Per01, PR03b]. Machine [PI05, Per07b, Per09b, Per11b, Per12b, PBRB08, PBS08, QCDM06, RBF06, Ric02, Ste95, SWSY11, WYSS10, dRF01, vW97, Bee89]. Machines [BBBB08, DFG^{+}12, FGK^{+}10, GKOT00, LV02, PLMV12, SW03a, WK08, ZT04, BF89, BHR84, BGR03, Sip97, Sze94, DK89, DLM07, MR01, Mar05a, Has96, Roc92]. machining [Hel91]. MACOM [SBV^{+}11, VBS^{+}10]. Madeira [JJF03, LN04b, MPF95]. Madhu [Gol11b]. Madison [CR04a, Pag98]. MADNES [BY06]. Madonna [BCL03, MSW05]. Madras [Thi94]. Madrid [ABB^{+}11, ABFH09, AGCP^{+}11, BH10, BK04a, CRL01, CS07, CP10b, CMM01, DAB01, Des09, DGSN11, DJJ12, Dor02, EWZ11, FKP^{+}05, Fer08, FOGE^{+}09, FR99a, FG04b, GMS03, GN12c, HP94, HP02b, HdMKS12, KOS02, LG99, Leu03, LA94, MPP08, Mar11c, MSS03a, MAA91, PA03a, PvES12, PP12b, PTZ03, dFEN04]. Madurai [KP06]. Maebashi [TYNM05, Zho01, ZRTS03]. Magdeburg [MK03a, OD12a, OD12b]. Mai [HdMKS12, Mah04, TKSN11]. Main [CE95, DZ10a, GSW06, MF11c, PPCS12, Tou96]. mainstream [GZR02]. maintenance [Kra97, MR91, Smi98]. Maiori [DY06a].

Making [HSL^{+}99, Mah04, THL11, Bui87, WPR08]. Malacca [HK10a]. Malaga [BW10a, DF01b, DGM09, LQO04, Mag02, MS95, TG10, dBBC02, dRWK^{+}10, BBC^{+}08, BHV11, CRJ11a, CRJ11b, DAC^{+}12, FG10, GVB12, HM12a, HC11, JL08, LH08, MRC12, MSK^{+}12, OAdGBP00, Vit10, YAKMJ11, HM02, WTM^{+}02]. Malaga-Torremolinos [MS95]. Malavella [EO94]. Malaysia [AIL12, CWYC12, CFL10, DV05, GG07a, GG07b, GG07c, HS00, HWG10, HK10a, Kur07, LM09, LLL06, LS09, MZ99, RWSS12, RK12, SZC^{+}03, TCHB12b, TCHB12a, ZRP^{+}09, ZRP^{+}11a, ZRP^{+}11b]. Malente [BMS04a, dRLP98]. Mallorca [BR97, BL05c, CGBDA08, EP06, LS04b, LSF07, Luo04, Luo05, Luo06, Luo08, Luo10, MC05, NL01, PH02, PCPS03, PD04, PF06, PF08, PF10, PFM12, PSS98, TJ08, VY93]. MALLOW [DFPV11, PAV^{+}10, SPON08]. Malmö [LN03b]. Malo [BN09a, BHH04b, BHH04c, BC92, CJ08a, CD07, CCM08, DL04a, FS07d, FCDB07, GANACBR10, GQ95, HAB98, KJB09, MR97a, MNW98, OS04, SRT06, VMvN12, VZM^{+}10, Coi99]. Malta [BFF^{+}10, EM05, MC94]. Malvern [APSZ02]. Malware [BL06d, FB09, HS07, HB11, KJ10, Zam08, JK05a]. Mammography [ABRZ06, Kru08, MOFM10]. MAMUS [KCO04]. Man [PS11a, Cea76a, Cea76b, DEM81]. man-computer [DEM81]. Managed [KMF06]. Management [AK10, AS11a, AB12, APS08, ASP01, AH02, ADR07, ACK00, ADM^{+}12, ABD03, ABT04, AH07, ALZ11, BB07, BGG^{+}12, BGK12, BG09, BBGR03, BG11a, Ber02a, BW10b, BF06, BMYZ91, BMYZ95, BK04b, BH06d, CG12, CLS11, CWC^{+}07, CUDL02, Che99a, CLZ^{+}09, Che10, CTYG10, CP06b, CCBW11, CP10c, CGS07, CK07, CCSW12, CLBP11, CFB11, DH05, DEKR09, DKK08, DPW04, DH10a, DTW03, DLW^{+}07, DPNM12, DRS08, DFS06, ED06a, FWY05, FKB02, FOGE^{+}09, FX08, GLW^{+}12, GANACBR10, GANACL11, GANACBdCdV12, GL10b, GP09b, GZ09, GPB02, GL08, Ham08, HT09, HMT10a, HT11, HKW11a, HHMT12, HL05, HS08b, HJMP06, HhKRT06, HTMC09, HMT10c, HKH10, HLFS12, JPD04, JP05b, JP06, JP07b, JP08, JP09b, JP10, JP11b, JP12a, Kam99, KR10a, KA12, KL07a, KR02a, KR04a, KJ09]. Management [KT06, KPR07, LLW06, LWF04, LFP^{+}09, LM08b, MCA08, MFG12, MNPW07, Mel04, MSW02, MKB^{+}09, NT03, NSL09, PMÓ06, PAD08b, PB09, PS07b, PF07b, Rei03, RK06, Ria08, Ria09, RK12, RMTW11, SPG10, SP09, SNČ^{+}12a, SNY06, SW04, SS09a, SS11c, SS05b, SJDT08, SVS08, SHW09, SZC^{+}03, SPÖ^{+}10, SF10, SS07c, SS07b, SS09c, SS11d, SLSW12, SvdMOP06, SD10, SWMM06, UWB^{+}06, VS09, VY93, VDD^{+}11, WLO^{+}11, WJL^{+}12, WM04, Wim03, Wim04, Win09, XL11, YBM87, Yam08, YKL06, ZS07, tTVH^{+}12, tHBP08, vEDA04, vdADO00, ALM95, ADB^{+}05, Bak95, Bar03, BC92, BB96, Bha95, BJS99, Che03b, Con97, DC00, Est95, Est99, Fei03, FS99a, FNY93, GSV05, HIS05, JLY03, JKB04, JP04b, Koh06a, LZ00, MMD05, Mag98, MXZ05, Mic90, Per05, Pie05, PB97, RKS89, San02]. management [Sch96, Sch98a, SPZ90, SHGB05, SXC04, Som96, SS99, TLF01, WYL01, WvdH03, vdAHW03, vdABCC05, AH02, Cho02a, DH05, FKB02, Hal96, HTMC09, KT06, MA03a, PAD08b, Per02a, VH04, vM94, Gun88b, Mat91]. Managing [PPPM10, Wes99, MD01, AH07, BB96, BK04b, CCBW11, CGS07, DKK08, MDY07, SS06c]. Manaus [NdMO03]. Manchester [ABRZ06, CS97, FGB11, Fle99, GK95, GMS12, LB05, Lou92, Lou94, ONC^{+}11, PMH09, Pit89, Sak01, Sch06, YAF^{+}02]. Manfred [ELS^{+}10]. Manga [MÁ07a, MÁ07b]. Mangalia [GNTF02].

Mangaratiba [DMB^{+}07]. Manhattan [Mel86]. Manila [HO98]. Manipulation [Ng79]. Manna [Der04]. Manno [Wes01]. Manual [BM03b, Mos04, TDB^{+}06, Uni81, JW74, JW75, Lis81, Luc87, TD97, TDBP01, UA83]. Manufacturing [MMV04, MVC07, MSZ09, MVL11, ALM95, MPB05, RD84]. Maó [MA03b, MA03c]. Map [MP06]. Maps [LH11, MP06, MSG07, MG08, PM09]. Mar [MÁ07a, MÁ07b]. Maranhão [BL04a]. Marburg [HLFS12]. March [AL11a, AI06, AJRB09, AM03, AF02, AH03, ANR99, AP04a, Ama08a, ABG96, AL11b, AB81, AMM99, AP83, BV05a, BS11a, BTT90, BDZ04, BH06b, BBH^{+}05, BGH06a, Bar11, BEM91, BWAM09, BL05b, BD97a, BDd09, BMSHW09, BF11, BCV06, BCP^{+}04, BCP99, Bet83, BG93b, Bha05, BN91a, BCB97, Bir12, Bir07, Bjo87, BV12, BMS04b, BZ83, Bla03b, BDC12, BGPW05, Boh75, BBD97, BGP00, BGR03, BRM06, BBPP09, Bra99, Bra07, BBK^{+}11, Bro92, BKM08b, BSL^{+}09, BKB03, BMP96, CJ11, CG12, CKD11, Can08, Can12, Cas09a, CFGN11, CSM10, CCL02, CUDL02, CW09, CCAS12, CCMR09, CCMS11, CSN14, Cle99, CAZ05, CSA^{+}12a, CSA^{+}12b, CHK06, Cra08, Cra12, CGvR02, DDS12, Dan12, DN88, De 07b, DMV12, DV09, DG12, Deu81, DSV01, DR06, DO89a, DO89b, DH04]. March [Din03, DMB^{+}07, Dov05, Dro08, DKR02, Due04, Dun12, Dur12, DL07, Dyc94, DHSH96, Ehr85a, Ehr85b, EEK^{+}87, EKLM87, EKR91, EW05, EM10, EFZKM07, ECV^{+}10, EEM^{+}11, Est98, FR98, FMO05, FI08, Fin99, FS99b, FK12a, FLF^{+}08, FZ86, FM96, GP07, Gan88, GW96, Gel09, Gel10, Gel12a, Gel12b, GBC^{+}08, GO11, Gib12, Gor10, GS00b, GSW06, GT05, GH04, GH07b, GHI^{+}06, GHP^{+}08, Gun94, Gup10, GKOT00, GHK^{+}10, Gut94, HR06, Han98, HKP08, HBS00, Hen08, HRB12, HP06, HAB98, Her91, HT06a, HW07, Hof11, HMSU04, Ibr95, IZ99, ISS^{+}06, Ish11, Jäh99, JT09, JBJ94, JJP^{+}02, JP04a, JPD04, JRR03, Kar91, KTC^{+}05, Ker97, Ker12, KEM07, KSB10, KR00a, Kit88, KHG01, Kno11, Knu99a, KKM^{+}11, Kos98, KP09, KO07b]. March [KU01, KH12, KBD87, KS10c, LLWL04, LBCO04, LO98, Leu08, LMA09, LMT^{+}04, LSK08, LCY03, LST^{+}12, LFL05, LS06b, LSU07, LS12b, LTT07, LV07b, LK00, MOP^{+}08, Mai00, Mai90, Mal97a, MM08, MS96a, MVOF08, MMT97, MKT98, MP95a, MT99a, MT86, Mit05, MP12, MT05b, MG07b, Mos97, MCER10, MSUB04, MNW98, MZ06b, NJHJ08, NLŚ10a, NLŚ10b, Nie94, Nip98, Niv98, O’B12, OVG^{+}08, ODP^{+}07, OZ01, Ölv10, Ong10, PCN12a, PCN12b, PCN12c, PY05a, PS12a, PT08a, PSD12, PMD90, Pie10, PRdRW09, PDG92, Por98, PQ05, Pup99, Rah04, RG05, Ram01, RR08, RD12a, Rei09, RM97b, RW86, Rob06, Rol98, REPH06, RT10, Sah13, SYNC11, SSRA98, ST08, SF74, SCM88, Sch04b, Sch12, Sei07, Sei12, SHGB05, Ses06, Sil01, SMK01, SMEGA10, Smo00, ST07, Som96]. March [SR11, Ste98, SW75, SFV^{+}11, STLK05, Swi99, TR12, TFK11, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, TCYZ12, Tho99, Tiu00, TO08, TG02, TST09, TGM97, THT03, VvS99, Val06, Vau98, Vck99, VDD^{+}11, VW08, Vit95, Wal04b, Wat00, Wei79, WMS04, WCBD08, YAKM^{+}12, Ytr06, Zan00, ZTY^{+}05, Zit01, dA09, dLZ12, dMS09, dMGFS11, vEDA04, vHC08, vH83, vdL01, vM94]. March/April [GS00b, HAB98, Mai00, Niv98, Ste98, Wat00]. Marciana [FC93]. Mare [AMT03, AMNT06, AHJ07a, AHJ07b, AHJ07c, EFZKM07, EHMM09, FSV09, GG98, MT02a, MTV09]. Marek [ACÇ^{+}12]. Marettimo [Fuc10, RF12]. Margheria [MR97b]. Margherita [Arm02, DPP95, GT04, RTA95, San92]. Marianske [Bec75, SPK01]. Maribor [D’H10b, ERW99, GKMV12]. Marina [FC93, GYA^{+}06, Ish97, KSHM09, PISW05]. market [LPH00, Zui99]. Marketing [Per08, Per06]. Marketplace [LPSW04, LY01]. Markets [FS07a, Vei03]. Markow [Her02]. Marktoberdorf [BS76]. Markup [AST05, AB04, Koh06b, SW03b]. Marne [BCP99]. Marne-la-Vallée [BCP99]. Marrakech [HMG03]. Marseille [AL03b, APM08, BBH^{+}05, Cal82, CBC^{+}02, DT06b, FTZ08a, FTZ08b, FTZ08c, FTZ08d, KS91, LN04a, Ng79]. Marseilles [LCF^{+}04]. Marstrand [Jeu98]. MARTES [Bru06b]. Martigny [BB05a]. Martin [LLM09]. Martinal [GI03]. Martinique [PH04]. Martino [CW94a]. Maryland [BKA11, FNY93, Lam96, TDS94]. Marz [SF74]. MASN [CPGL05]. Mass [KCO04, PS07a, PS08a]. Masses [CP10c]. MASSIVE [Lin01]. massively [Bre97, IGN05, JSS08]. MASTA [NSM07]. mastering [MSS12]. MATA [KKM04, MKPV05, Hor00, HMG03, KMD02, PG01]. Matching [AP10, AT02b, FL08, GM11a, KS12a, KU09, LV06, MZ07a, SMD04, AL01, Apo92, Apo93, AH97a, ACP05, BYCC03, CG94, FC98, GU95, GS00a, HM96, PC99, Vos92]. matchmaking [Vei03, Vei03]. MATES [BLKP08, BvdHPP09, DW10, EKL^{+}05, FTAZ06, KO11, LDTU04, PMKG07, TG12, SKMT04]. Mathematical [Bis88, CPRS01, DT89, FvdHJT10, GPS03, GKS05, IT06, KKMW07, Koh06b, KS10a, KS12c, KMS12, LB90, SKK04, TDG91, CC93, CC95, CCP96, HS05b, MM82, van90, ABH12, ASLM09, ABD03, ABT04, Bec75, Bec79, BJ86a, Bli75, BF06, BS93, Bro92, Bro94, CGMQ08, CK84, CJK88, DFL^{+}10, Dem80, DR02b, Ehr85b, FKK04, FvdHJT10, Gru77, GC81, GRW86, GI03, HK92, HK10b, IT06, JS05, Koh06a, KBČ^{+}12, KU06, KN09b, KMS89, KK07, KPW99, Mai88, Mai90, Maz76, Mel86, MS11b, NR00, OT08, PS96a, PRR94, PR97, Rov90, RV03, RSW12a, SPK01, SPO11, TPB79, Tar91, WH95, WR09, Win78, Zil87, ZBG98]. Mathematics [Alt01, ACR^{+}08, ACD^{+}10, BW03a, CDV03, CDCW09, DFUR11, EJM10, GS01, JCC^{+}12, KKKS11, Kap08, WDPW02, WDPW04, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, BCG02, BV92, EK81a, Nic75, Nic86, AAA^{+}11, APM08, BO00a, BMW93, BM02a, BDK10b, GN12c, HMS09, Jeu98, Koz04, MBS05a, MSW07, Mol95, Uus06, Van89, WM03]. Mathematische [BHR84]. matrices [Pan84, OZ83]. Matrix [GBDM77, Tha84, Smi74, SBD^{+}76]. Matsushima [ODP^{+}07]. Maude [CDE^{+}07]. Maurer [CRS11]. May [AP09, Abr01, AB11, AYB95, Ada96, AH94, ABB^{+}11, ACL12, ASE^{+}07, AJM87, ADJ^{+}01a, ADJ^{+}01b, AvAS06, AvASD06a, AvASD06b, AvASD06c, ANS^{+}09a, ANS^{+}09b, AS06, ASR^{+}07, AMRŚ08, ABS91, And96, ACL98, APS06, AGS^{+}11b, AGS^{+}11a, AT07, APN08, Apo92, Arn90, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, ALZ11, BL05a, BCB99, BC95, BE06, BSvRW08, BSvRW09, BBvRL10, BR03, BW11, BBS07, BN94, BBH^{+}10, BM05b, Bat09, BHP95, BS06a, BBK12, BBDS12, BS06b, BJP12, BCC^{+}10, BDG^{+}12b, BCD04, BMS04a, Ber08, BH07, BL10a, BR97, BKL^{+}12a, BKL^{+}12b, Bih03, BMSW12, BPS06, BMM07, BL89, BS94, Bv97, BDDS07, BHR84, Bor94, BGPR10, BGJ11, BAP^{+}05, BR07b, BR09, Bra75, BBT10, BCKT05, BCF06, Bri95, BJ12b, BtHRBB12, BHJY07, BAWH00, BZ08b, BHM^{+}02, BDFS04, BPS02, BKO04, BFC^{+}09]. May [BL11, CC04, CCL06, yCCZ07, CFI06, CD10, CGL^{+}09, CBG^{+}10, CBS^{+}12, CHB^{+}12, CV08, CGMP11, CL07, CCJ07, Cha05b, CWYC12, CWiM^{+}09, CGL^{+}11, CLY02, CC09a, CWZ11b, CHG^{+}07, Chi79, CM09a, Cho88, CQSS09, CFG96, Chr91, CM06, CB06, CLB11, CV10, CPvdH03, CMM93, Coh96, CS02a, CAB10, CDN12, CKMS05, CJ95, Con97, CMV96, CR73a, CR73b, CS02b, CKW11a, CKW11b, CAdB01, CFP94, CCR86, CS01b, Cra05, CvRVSN12, CSSW04, CFRR10, CMS08b, CHR10, DSZ04, Dam91, DM06, DS05a, DF01a, DPLW08, DSRW08, DS09a, DW07a, DSB10, De 95, DFPV11, DHtM^{+}12, DFMV10, DIMV11, DFK91, DF11, DBB12a, Des09, DP10a, DGWB09, Die02, DvEF06, Dji89, DRS05, DGR92, DIP12, DPMP^{+}11a, DPMP^{+}11b, DLW06, Dwa00, Dwy01, EKF12, EdS08, Ekl94a, Ekl94b, EW04, Eng81, FBSMS08]. May [FK10, FS07a, FMC04, FM97, Feh11, FR00, FS99a, FSdlPT11a, FSdlPT11b, FR09, FFL^{+}12, Fes10, FPCK11, FBM12, FSN06, FKS10, FHP03, FS09b, FSTS09, Fri04, FHP11, Fri93, Fum97, FN12b, GP09a, GKL09, Gal97b, GJ09, GAL12, GCFA12, GS01, GVT08, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, GGK06a, GWS05, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, GCK06, Gil10, GT04, GGS11, GGHS07, GPE05, GSP01, GR06b, GCC^{+}02b, GCC^{+}02a, GN08, GP89, GDGB12, GQ95, GS94, Gün88a, HMRT05, HKR07, Ham00, HK94, HHLS08, HKS06, HCS^{+}05, Hel94, HIS05, HS95b, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, HK11, HPS09, HR08, HCL05, HW03, HM05a, HPZ10, HYPY10a, HYPY10b, HM11c, HdL12, HSL06, HKM^{+}07, HCS11a, HCS11b, HMBD09, HKS09, HS98d, IdKM06, Ima99, IPRO08, IST07, IJ06]. May [JSV11, JSS08, Jan88, JLST06, JLM00, Jen03, JM95, JFT11, Jon90, JDMB08, JDMB10, Jou85b, KMR^{+}05, Kat99, KL05, KIRK05, KD05, Knu02, KW07b, KNP01, KHSM^{+}08, KS05a, KPK^{+}10, KPS09, KFMM06, KI12, KHI06, KHK^{+}08, KVMH09, Koz82, KKGT03c, KKGT03a, KKGT03b, KŽ10b, KDWW10, LLT07, LRS03, LGK^{+}04c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, Lam96, LBP00, LKK^{+}07, LBP06a, LBP06b, LBP06c, LBP06d, LG99, Les87, LOS05, LBO05, LCB12, LM96, LN01b, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, LPR08, LK09b, LN04b, LP00, Lou92, LP08, LGS09, LO88, Mac06, MPV12, MRK05, MRvM07, MJ04, MZ07b, MSZ08, MNZ09, MS03a, MR01, MT02a, MSS07, MS05, MPR07, Mau89, Mau96, MRPM10, MRP11, MPR12, McC96, McG08, MZC06, Mei08, MSS03a, MA11, MSH08, Mid01, MMMO96, MP93, MKR^{+}04, MMK^{+}05]. May [MS08b, MF06, MNS95, MPS04, MBC^{+}97, NMPM08, Nao07, Nes95, NGHJ07, NP10, Nic75, Nik05, NT03, NS99, NVSB^{+}07, Nyb98, O’H98, ONPR07, OT11, OSV11, Omi00, OSC^{+}08, OWM04, OYA04, OK96, PAV^{+}10, PZ07a, PHZ11, PM04a, PM04b, PM04c, PM04d, PTRH05, PT09, PP12a, POP09, PR11, Pat11, PRK12, PR05, PH97, PvES12, PP12b, PRT06, PT08b, PPS03, Pfi01, PMDA96, PMWO02, Pin89, Pod95, PVC^{+}99, Pol99a, Pol99b, PGV93, Pre00, PPF^{+}00, RPR08, RM09, RKM^{+}12, RTA95, RA00, RST09, RBF06, Reu99, RM04a, RYG11, Rol00, Rue93, Rus00, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, SSVW12, SHN07, SS83, San92, SMBQ07, SH12b, SJDT08, ST12a, SLH04, Sen10, SMP07, SOS96, SLB05, SCL^{+}06, SLL^{+}07, SvADS07a, SvADS07b, SvADS07c, SvADS07d, Sho84, SPON08, SCP^{+}12, SLSW12, SSM08]. May [SKK04, SSB90, Ste99, SBF08, SWMM06, SLP05, Stu90, SV01, SvASD05a, SvASD05b, SvASD05c, TCHB12b, TCHB12a, TG04, TC95, TT10, TDG91, TBF^{+}09, TNP^{+}07, TEB09, Tri98, TWSN07, UDF06, VJ11, VW07, Vau06, VHM07, V^{+}06, VSMS12, VP04, VASP09, VKG12, Wan03, WLY05a, WLY05b, WLY05c, WYZ06a, WYZ^{+}06b, WPRW06, WPR07, WPR08, WLGB^{+}08, WGMP09, WZH^{+}07, WSTI08, WvdH03, Wet92, WPM07, WG11, Wha03, Wii04, Wii05a, Wim03, Wim04, WWC02, WYX08, Yao82, YLW^{+}07, YKMJ09, YAKMJ10, YHZ09a, YHZ09b, YHZ09c, ZGK^{+}07, ZLY07b, ZJS^{+}02, ZT04, Zip92, dBdRR89, dBdRR90, dBdRR91, dBBC02, de 98, dS11a, vHH09, vdBHOS09, vdHCD07, vdHS10, vdPS08]. Maynooth [CFM09, MSS05b]. MB [DS11c]. MBC [CLLD09, CBS^{+}11, WZH^{+}12]. MBIA [LSIT11, YLS^{+}12]. MCAM [SLZH07]. MCBR [CMSM^{+}10, MGSM12]. MCBR-CDS [CMSM^{+}10, MGSM12]. MCD [RTZ08]. MCIS [CLLD09, YMY^{+}10]. McLean [THR03, Ada96]. MCM [AAA^{+}11]. McMaster [Lau93a]. MCPR [COMTLB12, MTCOK10, MTCOBH11]. MCS [BKR09, EKR10, HKR07, OPKR05, SKR11, KR00b, KR01, RK02, RKW04, WR03]. MCU [DLM07, Mar05a, MR01]. MCV [MLTC11]. MD [BKA11, BR12b, CSM10, DJ08, EFM04, HRTR05, HRR^{+}06, PT06b, RBF06, SYNC11, SBF08, YGE12, GM99, HRT^{+}03, Par02, Ras01]. MDA [AAR05, PS07a, PS08a, Leo00]. MDAFA [AAR05]. MDAI [TNM05, TNY07, TN08, TNI09, TND10, TNYL11, TN04, TNVDF06]. MDD [Bru06b]. MDDE [CUDL02]. MDGS [ADB^{+}12a, ADB^{+}12b]. MDIC [Tuc01]. MDM [Che03b, TLF01, ZSD03]. MDM/KDD [ZSD03]. MDSP [BGG^{+}12]. ME [EGMW11, AS99, EH95]. Meaning [ABdJS10, AKR^{+}12, Fre98, FJM06, YKMJ09]. Meaningful [MTP^{+}02, MTS^{+}03, MT^{+}03, MT04c, MT04d, MTC04, MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MTH08, MDH09a, MDH09b, MHD09, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, MT05a, MTH05]. means [Mul93]. Measurement [ABD^{+}09, BP04a, CU08, CGBDA08, DBB^{+}08c, KP10b, MTU09, NP11a, NP12a, SR11, TR12, UPB07, Dov05, DA01, MCER10, Sch12]. Measuring [BB95]. mechanics [Sac04]. mechanised [GMW79]. Mechanisms [FPRAW04, LSJ06, FRA05a, FH01, PSP^{+}02, MÁ05b, SD10].

Mechanized [KKMW07]. Mechanizing [HS05b]. MedAlg [ER12]. Medford [OSRS07, SCHS07]. MEDI [ABB12, BP11]. Media [ALPH10, AHSC11, ACHH12, AKSO06, BMN05, CKM^{+}09, DGG^{+}11a, DGSN11, DHH^{+}08, FSA^{+}07, HM12c, LYhL^{+}01, LWYN09, ZCGH11]. Mediated [PSP^{+}02, WBD08, DS01b, MSY00, NS99, FPRAW04, FS07a, LSJ06, PS09c, vEDA04, DC01, FRA05a, PSP^{+}02]. Medical [BHH04b, BHH04c, BS06b, BKLP11, BKLR12, CS05b, DK02a, DK02b, HSL^{+}99, KL07b, KLP10, LPZ12, MCKB06, MLTC11, OMMSP06, Per08, PPM09, SKK04, SWSY11, SH11, TN03, TGM97, WYSS10, BMSMS04, BG93a, CH91, DG97, IL01, KSTP99, AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, BH00, CMSM^{+}10, CGS02, CM04b, CMM01, DDJ00, DSL08, DG05b, DG05a, EP03a, EP03b, FMP11a, FMP11b, FMP11c, GML^{+}08a, JNPV10a, JNPV10b, JNPV10c, LNS06a, LNS06b, LEP^{+}10, MLTC11, MAFS08a, MAFS08b, MGSM12, NV01a, PBH03, TC99, WC98, YJ04, YJS^{+}06, YHR^{+}09a, YHR^{+}09b, Zha07, ZS10]. Medicine [BAHH07, CSAH09, DKB03, FN12a, Hol07, PLC11, Per02a, PS07a, PS08a, TGM97, Aya95, BSW95a, HSL^{+}99, Ker97, MHK05, Per05, Per06, QBA01]. Medina [Bar08a, DZGS06]. Mediterranean [ER12]. MeDoc [BBEdK98]. Meeting [AD98, Bar08a, BVAT12, CHD07, CR01a, CD01, CDND11, DK90, DKD05, DLM99, DKP00, DLO03, GY08, GSS09, KGRD10, KKDV02, KKD04, LKD08, MY06, MRU12, MTV09, MPT10, MTWD06, RL11a, RWD09, SBH^{+}04, TBD12, BDW97, RR01]. Meets [FMPT10, KKS11a, KKS11b, ZLY^{+}07a, Wei97]. MEGA [SLL^{+}06]. Mehlhorn [AAN09]. Melbourne [ATM09, BS02a, BSN06, BS01b, Che03b, CNA08, DDE04, Dig04, FPRAW04, GMO03, GHSW04, GDG11, HENR03, HGZ05, IZ00, KHJ05d, KRH09, Kow01, LOST04, LKZ^{+}08, MS00b, MSS05a, NL09, PH11a, Pie08, SYY^{+}08, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, WDKK07, WKK98, XCHZ11b, XCHZ11a, XLKZ12, Yao95, ZS00, ZS07, ZNCF03, vSH06]. Mellon [BRW85, KC84]. MELT [FK09]. Membrane [CFP^{+}09, EKP^{+}07, FPRS06, GHP^{+}11, GPR^{+}12, HPRS06, MVMP^{+}04, MPPJ^{+}05, Pau03, PPJRN^{+}10]. Memetic [PDSD10, PSDS11a, PSDS11b, PDSN12]. MEMICS [KBČ^{+}12]. Memorial [BBP96, BBZ00, BBZ01, BZ03, CVV12, PVV10, VV07]. Memory [CGMQ08, GRS06, MP10a, MSS03b, MCdS^{+}08, OKL04, Bod91, Cha05b, CKO01, EV01, Fre03, Vos03, Bak95, BC92, BMS96]. Memphis [DGSM11a, DGSM11b, GY08]. Menlo [SBM11]. Menor [MÁ07a, MÁ07b]. Menorca [AS06, HP89]. Mensura [ABD^{+}09, CGBDA08, DBB^{+}08c]. Mental [Bar02]. Merida [Coe02, BV12]. Merin [WTP^{+}06, VPBŠ03]. Mes [GB77]. Mesh [LPZ12]. MeshMed [LPZ12]. Message [CHD07, CDND11, DKP00, DLO03, KGRD10, KKDV02, KKD04, LKD08, MTWD06, RWD09, TBD12, AD98, BDW97, CD01, DKD05, DLM99, Koy92]. Meta [BJ78, FT94, BdBC01, FT94, Coi99, Pet92, Sch91]. META-92 [Pet92]. Meta-language [BJ78]. Meta-level [Coi99, Sch91]. meta-modeling [BdBC01]. Meta-programming [FT94]. Meta-programmming [Pet92]. Metaclasses [KS95]. Metaheuristics [AAB^{+}06, BBAB^{+}07, BBRS05, BBC^{+}08, BBD^{+}09, BBR^{+}10]. Metainformatics [Hic04, Nür03, Wii05b]. metalanguage [Bli87]. Metalevel [YM01]. Metaphors [Neh99]. MetaSoft [Bli87]. method [Ber98, Bjo87, BJ78, Lin01]. Methodological [HHL07b, HWLJ08b, BE05, Jac91]. Methodologies [CPZ07, DSLT08, DSLT10, DSHL11, BCBG05, BU94, BSA05, CS87, LPA00, MM82, PWZ98, ZL97, ABD^{+}12, PBG^{+}10, RZ91, RK93, RZ94, ZZ99]. Methodology [Ber02a, BF04, CD00a, DKW99, DD10, DPT08, ED06b, JV06, JP11a, KR02b, MR08, MdPA98, RA12, Sco02, WS06, AN95, Hac75, Hae99, HKB93, Joh97, LB78, Mar80, Rao95, RMS04, RB00, Rus00, Tob80, WP90, WN96, Ber99]. Methods [AO02, AG09, AM03, AdBB12, ABDW10, ACJ09, AT12, AGM03, BdB08, BPS11, BCK08, BMS04a, BMS08b, BJM09, BI03, BH07, BDZ08, BPZ09, BI11, BCP12, BVAT12, BHHJ11, BDS04, BJ07, BK04a, BDGN07, BKLS11, BKZ05, BNZ06, BB11, BC09b, BHLvdP07, BtHRBB12, BW09, BM11b, BKN07, BHLP07, BS11f, CG07, CGMQ08, CP12, CD09, CPV03, CF09b, CW03, CBBS09, CMS08b, DFL^{+}10, DSB04, DG07a, DSS11, DB04, DG09a, DGLT12, DDK11, DW03b, DZ10b, DH08, EF02, Ehr85a, EL02a, Fen00, Fis08, FHT05, FBHW00, GTP12, GT03, GKS^{+}07, GN12b, GM12, GO09, GW09, GL74a, GL74b, GP12, GW98, GPS03, GKS05, GW06a, GSS00, GI03, HU08, HS02, HBH08, HT06b, HJ00, IWC04, Jac09d, JMC10, JLW07a, KG12, KW01, KBČ^{+}12, KS10a, KS12c]. Methods [KR10b, KT10, LS10a, LK99, LM08a, LW09, LH06, LMA08, MWD06, Mar11a, Mar11b, MS06a, MS10b, MS10c, MS12b, MTV09, MPT10, MW03, MP03b, MM10, MA03b, MÁ07b, MNS06, NNS03, Neu04, NMP^{+}04, OW09, Oli07, OKL04, PS12a, PT91a, PT91b, Pri06, QQ11, RFZ03, RL11a, SS11b, Sch06, SMP07, SM11b, SS07c, SKK04, SP12, TMLL09, WW02, WNL05, Wol07, YJKK06, YJKSW10, YZCW07, dBBGdR03, dBBG04, dBBGdR06, dBBGdR07, dBBGdR08, dBBM09, dBBHL10, dS02, dS11a, dPMA98, ABL96, ASP85, BCG02, BHP95, Bec05, BF91, BC04a, BB05b, BB96, BJ86a, BN91a, BHL90, BBP93, BP05, BHK01, BJ95, BPS02, CE95, CB06, CKMS05, DS05a, DG01, DC00, Dim03, Dyc00, Ehr89, EHH12, Fer93, FP96, FZJ01, FJL97, FCF92, Gac02, Gal97b, GW96, GM03, God96]. methods [GSP01, GL03, HP99b, HM04, HSTU99, KT96, Kat99, KSW05, KMO^{+}05, KS97, LB05, Len90, MSS12, MM01, MMMO96, MP93, MT99c, Mur99, NDB94, OZ01, OZ83, PFCS05, PH97, Pri03a, RVY05, RSvdP05, Sif90, SC96, SZ05, VY93, WWD99a, WWD99b, WB89, WL93, ZEL04, dBBGdR05, de 98, BW91, BJRT09, EK77, HM11c, MA03c, MFÁ^{+}09b, Mül07a, RD84, VS02, VP04, Hus97]. Metrics [EFR08, KGS08]. Metrikon [DBB^{+}08c]. Metz [Kra05]. Mexican [ABG09a, BS11d, BS11e, CSCO00, COMTLB12, Coe02, GdATM05, GRG06, GM07a, GM08a, MTCOK10, MTCOBH11, MAFSS04, SAG10a, SAG10b]. Mexico [AB10, BYCC03, BS11d, BS11e, BCE09, BRU06a, BBR07, CSCO00, COMTLB12, CL10, Coe02, CAZ05, FK87, FMC04, FR00, FRL07, Gel01, Gel02, Gel03, Gel05, GdATM05, Gel06, GRG06, Gel07, GM07a, GM08a, Gel09, JRL09, KO07a, KCO04, LRG04, LO10, MTOK06, MTCOK10, MTCOBH11, MT08a, MT08b, MTH08, MCA^{+}07, MA05c, MAFSS04, Raj02, RUL04, RRU05, RCLE05, Rol00, STC04, SAG10a, SAG10b, STWS07, SC05, Tsu08, Whi00, WB09, ABG09a]. MEXT [FMY04, Oka03]. MEXT-NSF-JSPS [FMY04, Oka03]. MFCS [HK10b, JS05, KU06, KN09b, KK07, MS11b, OT08, RSW12a, DR02b, FKK04, NR00, PS96a, PR97, RV03, SPK01, WH95, ZBG98]. MFCS’93 [BS93]. MFCS’94 [PRR94]. MFCS’99 [KPW99]. MFDBS [Bis88, DT89, TDG91]. MI [KS08b, SS11a]. Miami [Baa03, CD05, Des02]. MIAR [LEP^{+}10, YJ04]. MICAI [BS11d, BS11e, SAG10a, SAG10b, ABG09a, CSCO00, Coe02, GdATM05, GRG06, GM07a, GM08a, MAFSS04]. MICCAI [AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, CPR^{+}10, CKP^{+}12, CMSM^{+}10, DG05b, DG05a, DFGN12, FMP11a, FMP11b, FMP11c, JNPV10a, JNPV10b, JNPV10c, LNS06a, LNS06b, LPZ12, LMCH12, LSIT11, MDY^{+}10, MDHB11, MLTC11, MAFS08a, MAFS08b, MGSM12, SWSY11, WYSS10, YHR^{+}09a, YHR^{+}09b, YLS^{+}12, YC11, YSL12, YHV12, BHH04b, BHH04c, DDJ00, DK02a, DK02b, EP03a, EP03b, NV01a, TC99, WC98]. Michael [BS11b, BDFM12, Len04]. Michel [BL92, HR95]. Michigan [SdDK91]. Michoacan [BYCC03]. Microcomputer [FHM82]. microprocessor [Hun94]. microprogrammed [Tha84]. Middleware [Alo05, BC09a, CC07, DHM99, GM10, IS08, Jac04, KK11e, vSH06, CPvdH03, ES03, GM05, Alo05, BC09a, CC07, ES03, Gue01, GM10, IS08, Jac04, KK11e, SC00, vSH06]. MIG [AF11, BCK10, EKO08, EGO09]. Mihir [Gol11b]. Milan [ABD04, BM05a, BW09, DRS08, FM97, FMZ05, HS95b, LPOA^{+}06, MPPJ^{+}05, ML11, STT11, dSG05, vF91]. Milano [Mar03]. Milestones [HHG11]. millennium [Knu99b]. Milovy [BSW95b, GP02, HJW00, JKB96, PTB99, PJ97]. Milton [GHK^{+}10]. Milwaukee [JSV11]. MIMI [GML^{+}08a, OBS12]. MINAR’98 [IS98]. mind [BND01, vdV84]. Minh [NHJ12a, NHJ12b]. Miniato [Bor93, PZ96]. minicomputers [Har77]. Minimization [BKLS11, CBBS09, RFZ03, YZCW07, FZJ01, HP99b, Naz94, PH97, RVY05]. Mining [ABG^{+}06, AGL^{+}07, ASR^{+}07, ACHH12, BHM^{+}04, BHMS07, BRM06, CGL^{+}09, CBG^{+}10, CFZ10, CZF10, CBS^{+}12, CHB^{+}12, CWiM^{+}09, CLY02, DSZ04, DKY06, DGDM^{+}11, EMT02b, GVB12, GSYZ10, GLC^{+}11a, GLC^{+}11b, GZSC07, HCL05, HYP^{+}09, HCS11a, HCS11b, Kam99, KMSS02, KŚHM11, LwCM00, LZhL06, LYT06, LLWL09, MM08, MLK04, MNLM06, NZS^{+}06, NSS^{+}07, NKLC06, Per02a, Per02b, PI05, Per07a, Per07b, Per08, Per09a, Per09b, Per10a, Per11a, Per11b, Per12a, Per12b, PRG09, PRG10, PRG11, Rei99, SCH^{+}09a, SLZH07, SBM08, SWS^{+}05, SYP^{+}05, TCHB12b, TCHB12a, TLZ^{+}08, TKCW11a, TKCW11b, TKCH09, TNA^{+}10, WGLL10, WZH^{+}07, WSTI08, WS06, ZSSM03, ZYRP10a, ZYRP10b, ZZ02, ZSM^{+}09, ZLY07b, ZKZ00, CWL01, GLS05, KZ97, LWD05, MMR07, PP99, Per01, PR03b, Per05, Per06, RH01, SXC04, TLC00, TYNM05, Wan03, W^{+}12, Wha03, WKK98, ZZ99, ZSO99]. mining [ZQ98, ŻR99, dRS01, RTZ08, ZSD03, ZH00]. Minneapolis [Li98, NS99, PTM^{+}11, WPR07]. Minnesota [Li98]. Minorca [MA03b, MA03c]. MIOS [MTH05, MTH06a, MTH06b]. MIOS-CIAO [MTH06a, MTH06b]. MIPS [RR04b, VC03]. MIRAGE [GP07, GP09a, GP11]. MIS [CC05, Wii05b, Hic04, JOD98, Nür03]. Miscellanea [Gol11b]. MIT-JSME [SLF91]. MITACS [GAL12]. MIWAI [SAUL11]. Mixed [Shu09, Shu11a, Shu11b]. Miyazaki [hKA10, SINK08]. MKM [ACR^{+}08, ACD^{+}10, BF06, CDCW09, DFUR11, JCC^{+}12, KKMW07, Koh06a, ABD03, ABT04]. ML [BvLR04]. MLCW [QCDM06]. MLDM [PI05, Per07b, Per09b, Per11b, Per12b, Per01, PR03b]. MLDM’99 [PP99]. MLMI [BB05a, BR06, PBRB08, PBS08, RBF06, SWSY11, WYSS10]. Mlýn [BFG^{+}12, NKM^{+}09, vLMP^{+}10]. MMAS [IGN05, JSS08]. MMB [Sch12, BB95]. MMB&DFT [MCER10]. MMBIA [SKK04]. MMCP [ABH12]. MMCS [DFL^{+}10]. MMI [LK09a]. MMIXware [Knu99b]. MMM [BTZ^{+}10, CCD^{+}06a, CCD^{+}06b, GKS05, HSMPA09, KS10a, KS12c, LTL^{+}11a, LTL^{+}11b, SNE08, SMH^{+}12, GSP01, GPS03]. MMM-ACNS [GKS05, KS10a, KS12c, GSP01, GPS03]. MMNS [ASP01, DH05, HJMP06, KPR07, PAD08b, PB09, VH04, AH02, MA03a]. MN [NS99, PTM^{+}11, WPR07]. MO [EM08, HCS^{+}05, IJ06]. Mobile [AZ11, BPS06, BAP^{+}05, BKM06, BY06, Chi03, CSRS08, CDM04, ENO12, FLR11, GCC^{+}02b, HJMP06, Jac11c, KG04, Kam99, KO07a, KPR07, KR06, LPR12, LK07, LK09b, LZ05, MKR^{+}04, NBK04, NW10, PAD08b, PB09, PBK03, RGLA09, SMBQ07, Sch10a, BDGM05, BB05b, DB02, Kim05a, KRMM^{+}02, KM00, Pat02, SC05, VJ99, Bau00, BD04, CSJD06, Che03b, CDM04, FK09, Gun94, Hor00, HMG03, JWH06, JMC10, KMD02, LhK03, Leo00, LK07, LK09b, Omi00, Pic01, PG01, RPZ97, RH98, Sur02, TLF01, Vig98, VT97, ZOCJ07]. MobileHCI [BD04]. MobileResponse [LK09b]. Mobiles [RBG09]. Mobility [CA08, GVCA07, IS12a, IS12b, KW98, KS05b, KKM04, MKPV05]. MobiMod [OYY03]. MOBY [BB05b]. MoChArt [EL07, PW09, vdMS11]. modal [FO95, PCG^{+}07, EV09]. MODAUI [Bru06b]. Model [BH10, BCP12, CV11, CZG^{+}08, CP07, Cou05, EN06, ELS^{+}10, GV08, HdL12, JS11a, JMO09, KOFJ09, KR12, LPZ08, Mel04, Pai09, PRR05, RR89, SL03, TG10, VGP08, Bee86, Cor02, Dam99b, Gog94, HPV00, JHC02, KS95, OK90, PD96, Sch93a, Zuc03, ABB12, AVP07, AAR05, BR03, BP11, BL02a, BE07, BW05b, BJK^{+}05, BKM08b, COB^{+}08, DP12, Dwy01, EL07, EOSW07, FKBA12, GKL^{+}11, God05a, GM04b, GM11b, HK05a, HMP08, Kon04, NWHR06, PHR09, Păs09, PW09, PRH10a, PRH10b, Qui04, RW06, SH08, SS09b, Val06, WCK11, vdPW10, vdMS11, Lev92, Van94]. Model-Based [BJK^{+}05, GKL^{+}11]. Model-Checking [Qui04]. Model-Driven [BCP12, ELS^{+}10, KOFJ09, BKM08b]. ModelAge [MS99a]. Modeling [ABH12, ADM11, ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12, ABH^{+}10, ABM05, AK12, AB06, ACR12, ADS09, AGCP^{+}11, AS07c, BBH^{+}10, BDW08, BDSW09, BTZ^{+}10, BCGY09, BCR03, Cai08, CJ11, CNP^{+}08, CJ08a, CVLL12, CSM10, CCD^{+}06a, CCD^{+}06b, CH10, CW07, CJ08b, CEFZM05, CW02, CMP07, DKC10, DMB^{+}11, DKM^{+}05, Duf07, Duf09, Duf11, ERK07, EOR06, EKS00, FT11, FRSH05, GLW07, GR06b, HRK^{+}07, HPTV05, HS02, HP09, HAB02b, HMPZ09, HSMPA09, JP03, JDL11, JHC02, JN12, KD11, KMC^{+}01, KS06, KCMO11, KH11c, KMVY06, KMYJ08, KJS01, LLS00, LCD^{+}09, LN04a, LTL^{+}11a, LTL^{+}11b, LFIM07, LLIH07, LSYO08, LFJI10a, LFJI10b, LJS^{+}10, MPK04, MMDN12, MA11, MÁ07a, MvL10, MSX10, MYS10, Nar01a, NABK12, OO12, OYY03, OPS11, OV09, PSST07, PSS^{+}10a, PY05b, PBG^{+}10, RT07, RRMP08, RBsC^{+}06]. Modeling [SS07a, SYNC11, SNE08, SMH^{+}12, SLLS03, SPC^{+}08, SPB00, SMK02, SWB03, SM09, TNP^{+}07, TNVDF06, TDK^{+}10, VWH06, VBH07, Vaz99, VAM06, WPRW06, Wii04, YGE12, YKMJ09, ZZTG07, vLS11, ABCWM99, ALS^{+}05, AW05, Ari02, ACL^{+}04, Bai05, BdBC01, BGV01, BM99b, Cea76a, Cea76b, CS87, Che99a, Che99b, CM03b, DG01, DTV02, EG97, FS99a, FR99b, GK01, JSdSA05, JM02, KMO^{+}05, LMT00, LRL98, Mag03, Nay95, PVV05, PV04, Rus05, Sac04, Tha96, WTZ^{+}04, Wii05a, YY01, Akm01, ACHH12, BCRB^{+}11, Bla03a, BSB^{+}99, Cas01, DKLT05, KD11, KMK10, KRRBV07, MSA99a, Rey04, RBSL06, TN04, TNM05, TNY07, TN08, TNI09, TND10, TNYL11, WK08, AD03]. Modelling [BSMM04, BB95, Bv97, CKP^{+}12, CW94a, CM95, DKS12, GB99, KS03c, KBC^{+}09, MCER10, PAC08, PM96, Sch12, THL11, VP96, Bit88, Enc80, Fie02, Hal96, HK94, HBS00, LY04, Lou94, MR97a, Pap95, PB97, PSS98, ZL96, ZGT05, BFvdM10, FKBP11, KSGT10, LCS91, LSR03, MTT98, MT09b, MA03b, SGD09, VTK^{+}12, vdMBD08]. MODELS [Cha09, DS11b, FKBA12, Gho10, Kie12, PRH10a, PRH10b, SS09b, WCK11, Bru06b, COB^{+}08, EOSW07, Gie08, Küh07, NWHR06, ADDL08, ACL12, AMZ09, ABF^{+}07, ANF08, ABR09, BdBC01, BEM91, BPR08, BV11, BGS^{+}03, BJ12a, BW05b, BJRT09, CCL06, CDP02, CPR^{+}10, CKP^{+}12, CSCR08, CZZ08, CC09a, CW06, CA10, CLS07a, DVVA08, DR11, DFM04, DDM08, DT06a, DKOZ05b, EPVW10, FV09, FP08, FHP11, FN12b, GPS03, GKS05, HT06b, JvdAB08, JvdA09, JBK09, JDK10, JDK12, JvdAM^{+}12, KMPH09, KY07b, KPK^{+}10, KS10a, KS12c, KGS08, KH11c, KLFK10, KS10b, LZ08, Leo04, LS05, LSW12, MP01a, MÁ05b, MFÁ^{+}09b, MV07, OT11, PPCS12, PR00, PPdV12, SST00, Sir12, VA07, Vit10, WJP07, Wol07, WMB10a, ZZ02, vdADO00, A^{+}08, AT02c, ADL09, BE95, BC07, BRR87a, BRR87b, BKS96, yCCZ07]. models [CNY95, CH96, CW99, DC00, GL97, GSP01, JP05a, MS99a, Mue01, San80b, Sch03c, SCKR03, Zil87, dBdRR89, dBdRR90, Cha09, DS11b, Gho10, Gie08, KW01, Kie12, Küh07, LNSS11, RRMP08, RR98b, Wes99]. Modena [KZ01]. Modern [IJ06, Dam99a, EK81a, Bra88, MMO^{+}11a, MMO^{+}11b]. MoDeVA [Bru06b]. MoDIC [CVLL12]. Modified [Mei89]. Modular [BS03b, GW00, LS06a, Mos97, NY83, BS03b, Ger04, GW00, LS06a, Mos97, MO97, Mül02, SPS09, Vog92]. Modularization [SPS09]. MOL [EJM10, KKKS11]. Moldova [GMV06, MR01]. Molecular [AGI^{+}06, BS11a, Bat09, Ber10, CPRS01, Cho12, DS09b, JPR04, MMK^{+}05, SM11a, SH07, ST12c, VW08, CS11]. Monastery [ABB99, KK11d]. MONET [MTH06a, MTH06b, MTH07a, MTH07b, MTH08, MHD09, MDH10c, MDH11a]. Monica [ŞS05a]. Monitoring [DPSU11, POP09, PSD12, RMB10, Smi98]. monotonic [DPP95, DPP97, Rei89]. Mont [BL92, BFGM05, HR95, PR05]. Montana [BCB97]. Montanari [DDM08]. Monte [GKOT00, JS08, Nie94]. Montego [BW05b, Bru06b, SV05]. Montenegro [ABMS06]. Monterey [CJ11, CG12, CS10, KS07a, KS07b, PM08]. Monterrey [GdATM05, MT08a, MT08b, MTH08]. Montevideo [BV05a, BPS11]. Montpellier [AF97, BPR02, BL08, Ber98, BG12b, BB02, DH04, FvdTDG12, GS01, GZR02, MT06a, MT06b, MTH06a, MTH06b, MD96, MC98, PR08, PH10, PP11, VAM06, Wii04]. Montreal [AGB^{+}11, BBD^{+}10, BY07, BEM06, GV09, KC07, KD05, KW07b, MMDN12, MO08, MMT08, Sch09b, Smi12, SRL12b, WANL08, AN95, BRP09, CFP94, Deh93, EP03a, EP03b, FKB02, FGM92, FGL96, GFV00, GS00a, Ham00, Har01, JY14, KO91, Kir93, KKGT03c, KKGT03a, KKGT03b, MKPV05, MB98, PU04, PG01, PBK03, RM97a, SRL12a, SRL12c, Tah00, WS96, Whi98, WMT96, WWC02, vP93]. Montreux [BRR97, Str96]. Mook [dBdRR90, dB92]. Moravicí [FKM09]. MORE-BI [CVLL12, DMB^{+}11]. Morelia [BYCC03, KO07a]. Morella [BBR07]. Morocco [EML^{+}12, HMG03, KS10c, MV12b, Vau08]. Morphology [KMS12, SPO11, WR09, AKR03, Ker01]. Morro [Mon01]. MOS [VT97]. Moscow [BUG93, BGU95, CW94b, HRSS08, KMKP11, KŚHM11, LKZCS05, RDK09]. MOSIX [BGW93]. Mosses [Pal09]. MOST [HP09]. MOST-ONISW [HP09]. Mota [Bar08a]. Motes [RBG09]. Motion [ERK07, JMBS07, MTT98, PH02, PD04, PF06, PF08, PF10, PFM12, Rüt98, CRYS09, Mac06, NL01, dPS95, AF11, BCK10, EKO08, EGO09]. Motivated [BLPW02, LBP00]. Mountain [JYZ04, STL11]. Mountains [WZJX12]. Mousson [Gal97b]. Move [MTP^{+}02, MTS^{+}03, MT^{+}03, MT04c, MT04d, MTC04, MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MTH08, MDH09a, MDH09b, MHD09, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, MT05a, MTH05]. MOVEP [Cas01]. movie [LZ02]. Moving [Wes01, Kat92]. MOZ [Van05]. Mozart [Van05]. Mozart/OZ [Van05]. MP/Model [RR89]. MPC [APM08, BDK10b, GN12c, Uus06, BO00a, BM02a, Jeu98, Koz04, Mol95]. MPI [CHD07, CDND11, DKD05, KGRD10, KKD04, LKD08, MTWD06, RWD09, TBD12, AD98, BDW97, CD01, DLM99, DKP00, DLO03, KKDV02]. MRC [RBSL06]. MRCAS [TGM97]. MRCS [GJTS06]. MSA [ZSY^{+}06]. MSC [She03]. MSEPT [PP12a]. Msida [EM05]. MSM [AHSC11, ACHH12]. MSN [CSJD06, JWH06, ZOCJ07]. Mt [HSB^{+}10, WZ03]. MTiP [Bru06b]. MTPP [HM11c]. Muggelsee [Bla90]. Mulhouse [BM99b]. Multi [BPD06, Bv97, BDDS05, BDDS06, BDDS07, BBT10, BJ01c, BHH08, BLVV07, CW94a, CM95, CZG^{+}08, CGG^{+}07, CDN12, CS12c, DLFP02, DDE04, DSLT08, DSRW08, DSLT10, DSHL11, DBB12a, DLW12, DFN10, DLGJ10, EFG^{+}09, FST09, FvdTDG12, GLZ^{+}03, GCL^{+}06, GB99, GGS09, GZSC07, HNH^{+}09, Han02, HPS09, IST07, Ish99b, IGN05, JSS08, JNHJ10a, JNHJ10b, JKN^{+}12, Kou11, KL02, LB03, LT05, LTÅ^{+}11, Lin03, LGR^{+}04, LS09, MMP03, MMV04, MVC07, MSZ09, MVL11, MPR07, MRPM10, MRP11, MPR12, MB09, Nag01, NcC99, NGHJ07, NJHJ08, ONC^{+}11, ODP^{+}07, PMMR06, PM96, PCG^{+}07, PVP05, RPR08, RM09, RWSS12, SS08a, Sch01b, SS06b, TDWG11, TGH06, TT06b, TtHVS06, TNGK08, VP96, VASP09, WPM06, WPM07, WBD08, YYI^{+}09, YY01, ZS00, AKK03, BK05a, CGLR05, CAZ05, DLT05, DL04b, DKN02]. multi [Fon03, Hal96, HWP01, KKA05, MPB05, MSS05b, PWZ98, RMR05, WR01, WS96, Wei97, WPM05, Zit01, AT07, APN08, Bv97, BGJ11, BLVV07, CKTW12, DS09a, DM01b, DLT05, DP10a, GB99, HENR03, KHG01, LMŠT01, LS09, MŠKL02, MMP03, PVP05, SCG98, SBD03, SA06, SAUL11, VSMS12, ZL97, ZL98]. Multi-Agent [BPD06, Bv97, BDDS05, BDDS06, BDDS07, BBT10, BHH08, BLVV07, CW94a, CM95, CGG^{+}07, CDN12, DLFP02, DDE04, DSLT08, DSRW08, DSLT10, DSHL11, DBB12a, DLW12, DFN10, DLGJ10, FST09, FvdTDG12, GLZ^{+}03, GCL^{+}06, GB99, GGS09, HNH^{+}09, Han02, HPS09, IST07, IGN05, JSS08, JNHJ10a, JNHJ10b, JKN^{+}12, KL02, LB03, LT05, LTÅ^{+}11, LGR^{+}04, MMP03, MMV04, MVC07, MSZ09, MVL11, MPR07, MRPM10, MRP11, MPR12, MB09, Nag01, NGHJ07, NJHJ08, ONC^{+}11, PMMR06, PM96, RPR08, RM09, RWSS12, SS08a, Sch01b, SS06b, TGH06, TT06b, TtHVS06, TNGK08, VP96, VASP09, WPM06, WPM07, WBD08, YYI^{+}09, BJ01c, CS12c, PVP05, AKK03, BK05a, CGLR05, DL04b, DKN02, HWP01, KKA05, MPB05, MSS05b, PWZ98, RMR05, WR01, WS96, Wei97, WPM05, BLVV07, CKTW12, LS09, Bv97, DLT05, GB99]. Multi-agent [LMŠT01, MŠKL02, MMP03, PVP05, SCG98, ZL97, ZL98]. multi-agent-based [DLT05, AT07, APN08, BGJ11, DS09a, DP10a, HENR03, SA06, VSMS12, DM01b, SBD03]. Multi-Agents [BK05a, BHH08, GGS09, GZSC07, Ish99b, LB03, LS09, NcC99, SS06b, YY01, ZS00]. Multi-Core [CZG^{+}08]. Multi-Criterion [EFG^{+}09, ODP^{+}07, TDWG11, CAZ05, Fon03, Zit01]. Multi-disciplinary [SAUL11]. multi-domain [Hal96].

Multi-image [KHG01]. Multi-modal [PCG^{+}07]. Multi-Party [Lin03]. Multi-Service [Kou11]. Multiagent [BMU^{+}09, BPD06, Hug03, Lin01, NKC09, Ish94, Kra97, BLKP08, BvdHPP09, DW10, EKL^{+}05, FTAZ06, Ish99b, KO11, KCO04, LDTU04, PMKG07, SKMT04, Sin94, TG12]. Multicomputers [HM11c]. Multicore [KKW11, KKW12, PP12a]. Multicore-Challenge [KKW11]. Multidisciplinary [CHR10, vEV12, Paz97, HRdV11, QBY^{+}12, SL12, VEG^{+}11]. Multilingual [PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PGBK04, AFP^{+}10, FGK^{+}11, PCG^{+}05b, PDK^{+}10, PCG^{+}10]. Multilinguality [CFH^{+}12]. Multimedia [ANS04a, ANS04b, ANS04c, AH02, AKSO06, BMO02, BTZ^{+}10, CCS04, CCD^{+}06a, CCD^{+}06b, CUDL02, CKM^{+}09, DSB10, DLNU11, DC12, DGG^{+}11a, DJNR06, DLN10, DGSN11, DKN^{+}11, DHM99, DHH^{+}08, FSA^{+}07, FLL03, HJMP06, HK05b, HXC^{+}08, HSMPA09, IAL^{+}07, JMC10, KPR07, LTL^{+}11a, LTL^{+}11b, LM06b, LXH^{+}12, LM03, ML04, MMBND07, MA03a, MZCC09, MWK^{+}09, ND04, PST11, PAD08b, Per02b, PCG^{+}10, PB09, QLK^{+}10, QLK^{+}11, RR04b, SÇS12, SNE08, SMH^{+}12, SvS00, Shi06, Shi07, Shi09, Shi10, SEK^{+}11, Shi12, SK12, Vaz99, VC03, WKD08, ZSD03, ZCCL10, ZYRH06, ASP01, BMP96, CC05, CCH02, DH05, DD97, DOS99, DKU05, GP98, HK05c, Hut96, JOD98, NK99, She01, SLC01, SW97, VH04, ZVU99, vSN01, BKS96, FM97, GJTS06, Hut94, HS98d, IS98, LG99, SLZH07, Ste94, Tuc01, VDPD96]. Multimedial [BDN08]. Multimodal [AFP^{+}10, ADM^{+}08, ECV^{+}10, EEM^{+}11, FOGE^{+}09, FGK^{+}11, Jac07b, PJM^{+}08, PDF^{+}09, PBRB08, PBS08, RBF06, TSG00, BB05a, BR06, BB01b, BBB98, EHMM09, KMPH09, LSIT11, SG07, SBF08, YLS^{+}12]. Multimodality [CFH^{+}12]. Multiobjective [BDMS08]. Multiparadigm [Van05].

Multiple [CDS09, Pol01, Bui87, Dam97, KvG98, BKR09, EKR10, HKR07, Ima99, KR00b, KR01, OPKR05, RK02, RKW04, SBV^{+}11, SKR11, VBS^{+}10, WR03]. multiple-agent [Dam97]. multiply [Pan84]. Multiservice [FCLG04, ABLM05, MB01, Mar03, RMV96]. Multiset [CPRS01]. Multivariate [BDFM12]. Mumbai [BHM03, Pal03, Sar03]. Mundici [ACG^{+}07]. Munich [BMP01, BNUW09, Bod91, BRW93, Bod96, BLKW01, DKM91, DL81, DBB^{+}08c, GWS05, GH94a, GH94b, GK06, GHI^{+}06, HRB12, ISS^{+}06, Jou94, KWA01, KM04, KFW04, LPH00, MRK05, MP95a, May01, RF01, Rig08, SNRG11, SRM^{+}11, SCKR03, WN96, WHR10, BBB^{+}85, BEH^{+}85]. Munster [BHR84, Dit88, JP09a, JFT11, RMFG06]. Munster/Westfalen [BHR84]. Murcia [RGLA09]. Murray [GF97]. MUSCLE [SÇS12]. MUSE [AHSC11, ACHH12]. Music [AAA^{+}11, DLN10, KMVY06, KMYJ08, MRC12, Wii04, YKMJ09, YAKMJ11, YAKM^{+}12, Tan93, Wii05a, AFS02, Lem97]. musicology [Lem97]. MWCN [Omi00]. My [GB77]. MYCRYPT [DV05]. Mylopoulos [BCGY09].

N [Del00, CCK95, KMN94, Tri98]. N.C [Mic90, RZ91]. N.Y [Ant97]. N’99 [Zui99]. Nîmes [ARS09, HLR^{+}98]. NAA [LVW05, MVW09, VWY01]. nad [FKM09]. NADA [MT99c]. Nafplion [Bra85]. Nagano [wLN99]. Nagasaki [EYX^{+}05, YAL^{+}05]. Nagl [ELS^{+}10]. Nagoya [Fur95, FU96, Iba92, RS99a, YYI^{+}09]. Nagoya-University [Fur95]. NAIS [AMNT06]. Naleczów [WDPW02]. Namur [JP05a, Le 94]. Nancy [BCG^{+}07, Bun94, BH06d, CCBW11, DN88, DH97, DRDKE11, Gan88, HMT10b, Jou85a, KW90, KR00a, LLP99, MM10, Mon96, Ret97, TC98, WHG07, vdABCC05]. Nanjing [CPGL05, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, PCG^{+}05a, PZE^{+}08, SLL^{+}07, WZH^{+}07, ZLY07b, ZW09]. Nantes [BD12, BJP12, Ben06, Cas01, EFG^{+}09, GO06, ST08, SC07, Tho06, ZQ98]. Napa [MT05a, MTH05, Sho84]. Naples [DDFM05, DPT12, Fes10, FBV11, GMP06, GBC^{+}08, GL08, MM08, MRSV07, NSS03, OVG^{+}08, SKR11, vHC08]. Napoca [HPSZ08]. Napoli [VC03]. Nara [Gie05, KS04a, KKKS11, LJS08, MKR07, OYM07, PT11, TBF^{+}09, Urz05]. NASA [Wil99, BHHJ11, GP12]. Nashville [DCFR12, EOSW07, Gie08]. Natal [CDGW10, CEM^{+}06, DSS11, GN12b, dCJB12, PSM10, YCB12]. National [BH06c, Bow94, BSW12, CK07, ENP02, Emb98, FGB11, GKL09, GK95, GL92, GJS08, GW93, JNS05a, JLY03, LJ00, Mac12, MK96, Rea01, Sex09, Sma97, WM04]. Nations [AM03]. native [Gor02]. NATO [GNTF02]. Natural [ABJ02, AMNT06, BDIR07a, BDIR07b, BIMW12, BKM01, BEKS02, DLŠ11a, DLŠ11b, DLJ12, FSdlPT11a, FSdlPT11b, Fuc10, GKKW09, HRM^{+}10, HMMW10, IK12, JWbG^{+}06, JWG^{+}06, KSS08, KLM^{+}07, KMPH09, KTB09, Kon04, KFMM06, KT10, KF12, LRH10, MM09a, MM04, MSA99a, MSA99b, MP01a, MP01b, MA03b, MA03c, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, MMM05, MR12b, MMM11, NR08, Paz03, RM02, RF12, SGPP06, VMBMN04, YJKK06, YJKSW10, AZ96, Dal92, Fra96, Hau86, Hir81, HVT12, MS95, Mit98, Pet99, RS96, SG96, WCO05a, WCO05b, WCO05c, ABJ02, BEP04, Bla88, Bol78, BIMW12, BKM01, Chr00b, DWSK05, Fil91, HRM^{+}10, HMMW10, KSS08, KLM^{+}07, KFMM06, MM04, MMM05, MMM11, Stu90, STLK05].

Nature [CCD^{+}12a, CCD^{+}12b, CR05b, GTP12, RJL^{+}08, RBB^{+}06, SCKR10a, SCKR10b, SKRZ04, DSM94, EV96, MAB^{+}02, SDR^{+}00, SM91, YBL^{+}04, MÁ07b, EBSS98]. Nature-Inspired [GTP12, SKRZ04]. nature-PPSN [EBSS98]. Naurod [May93]. navigation [Fre03]. NBiS [EBT07, TBE08]. NC [CGMP11, NQL06, ZCRL08, CPC^{+}99, RO00a]. NC-Pro [CGMP11]. NCUS [DsSL^{+}07, EYX^{+}05, ZSY^{+}06]. Near [JN01, Bli75]. Near-Optimal [JN01]. Needs [MKZ02, MKZB04, MKZK06, MKZK08, MKZK10a, MKZK10b, MKPZ12a, MKPZ12b, PM08]. Negation [Dau03, Nai86, Wag94]. Negotiation [FS07a]. negotiations [Vei03]. Neil [MSS02]. Nelson [BS06c]. Nepal [MAD^{+}04, NRR12, SU97]. Nerja [LT11b]. nested [Lev92, AFS89]. NET [AC09, CT07, NQR09, Apt00, Bra80, ERRW03, DB04, Bra80]. NET-COOP [AC09, CT07, NQR09]. Net.ObjectDays [WL04]. NET/FME [DB04]. Netherlands [AG01, AvLR91a, AvLR91b, ACvLdR03, ABD^{+}09, AFvZ02, ADP09, ABdJS10, AKR^{+}12, ACJ09, ABR09, Ame91b, AAR05, BK90, BG91a, BM99a, BLPW03, BBdB12, BEOT06, BSC12, BN94, BLS12, BAHH07, BCT06, BMSHW09, BCV06, BB01b, BG93b, BTPPS07, Bod03, BPD06, BLvdH10, BDDS06, Bos00a, BCK10, BIMW12, BL06c, Bri97, BHK01, BSH06, BAWH00, CPSA10, CD09, CCJ07, CP88, CW99, CA10, Cle99, CR01b, CLT05, CKTW12, DN04, DG10, DJS06, DvEF06, DTMH10, DS08, DvV96, EKO08, EGO09, ERRS10, EBSS98, EK10, Feh11, Fin99, FGK^{+}11, GFV09, GW03, GKST11, GO09, GH09, GM06a, Gue04, HM11b, HKB99, HP95, HZ10, HLMW06, Hee94, HW01, HS95a, HHW01, HB11, HPRS06, HS99, IdKM06, Jäh99, JSWhT11, KL07b, Kat92, KMS^{+}11, KHPS07, Knu02, KHL10, KBHSW06, KC00]. Netherlands [KPS09, KvdM00, KH12, LSJ06, LV07a, LAF^{+}10, LWRdR11, Mad92, MBS11, MEAC04, MdRIR08, Mid01, MvL10, MZ06a, MV06, MBH^{+}11, Nag90, NSM00, NdG04, NV01a, NQR09, NBM09, ORS89a, ORS89b, OFK05, PHR09, PMMR06, Pet03, PW03, PLG06, PBS08, PvS07, PT91a, PT91b, Pro96, Rau04b, Ria08, RSvdP05, RV98, RKNV09, SP09, SL11, Sch95, Sch04a, SvS00, SA06, SEL09, SBH98, Slo99, STDH02c, STDH02a, STDH02b, SCH04c, SHP^{+}05, SHS10, Swi99, TMdB11, TV94, Tho99, TGH06, TtHVS06, VvS99, Vad07, Van89, VP96, Vyt91, WPM06, WBW94, WR09, WJ95, dBNT87a, dBNT87b, dBdRR89, dBdRR90, dBdRR91, dB92, dBdRR93, dBdRR94, dBBGdR03, dBBG04, dBBGdR05, dBBGdR06, dBBGdR07, dBBGdR08, dBBHL10, de 98, dS02, dS11a, ter97, vB97, vEGW09, vES10, vEGSW11, vE91, vKS12]. Netherlands [vLdB80, vL88, vL89, vL94a, vL94b, vSN01, vdPW10, vdBHOS09, vdBK05, vdHP12, vdAB03, vdAHW03]. NetObjectDays [AMU03]. Nets [ADR01, AMT03, AMNT06, BR96a, CR04b, DRR04, DT06a, EJPR01, ÉEMS04, EL02b, FW09, HP12, JvdAB08, JvdA09, JBK09, JDK10, JDK12, JvdAM^{+}12, KY07b, KP11, LP10a, MT02a, MA03c, Roz93, vHV08, vdAB03, AB97, BDR99, BRR87a, BRR87b, CD05, CK01, DD95, DS98, Jen92, KD99, Mar93, NS00, RR98b, RR98c, Roc92, RRG85, Roz86, Roz87, Roz88, Roz90, Roz91, RRP91, Roz92, Tau89, Val94, ADR01, EJPR01, ERRW03, BDR99, CYR02, DRR04, Vog92]. Nets4Cars [SFV^{+}11, VMB^{+}12]. Nets4Cars/Nets4Trains [SFV^{+}11, VMB^{+}12]. Nets4Trains [SFV^{+}11, VMB^{+}12]. Network [APFV09, ABF^{+}07, AH07, BLO^{+}07, BSZ12, BP04a, BCL03, BGKY08, CKD11, CK12a, CA08, Cer07, aCF06, CWC^{+}07, CLZ^{+}09, CEK11, CC12, Cho02b, CU08, DG07b, FHW08, GMO09, GVCA07, GSYZ10, GPS03, HWG10, HBH08, HTMC09, IKY05, JYZ04, KY07a, KT06, KS10a, KS12c, LT01, Leo09, LTW11, LM03, LT11b, MCA08, MTU09, PZ08, PMC06, Sab10, SMK06, TNP^{+}07, UPB07, ZYH03, ZYB06, ZY10, dMS06, BB89, DD97, DY06b, DWML05, Dov05, GSP01, GKS05, KS05b, Kum90, PS97a, RMV96, SUW02, SLL^{+}06, SS99, TAS05, AC09, AS11b, BE05, CLWC08, CT07, DSZ10, EBT07, FL00, Her92b, JGXC04, JRJ05, Kou11, LJJG07, LG95, NQR09, PRV11, PS98a, PZYS12, Ran93, RST09, SBC^{+}94, SL99a, TBE08, XBM12]. network-based [PS97a, EBT07, TBE08, FL00, PS98a, SL99a]. Networked [ABD^{+}12, BI11, Bha00, Bha05, Bha07, BB03, BGKS04, CG04, DV07a, KSB10, KMSB11, KHW03, KS07b, NPPDG06, PV02, SCKR03, SHYEF08, Wan05a, WL04, dFEN04, AMU03, Bha02, Ish99a, AK10, CH01, RF99a]. NETWORKING [BBK12, CGMP11, AYV^{+}11, ASE^{+}07, ABK12, BMK08, BMK09, BDK10a, BKH11, BKL^{+}12a, BKL^{+}12b, BPS06, BDDM12, BAP^{+}05, CDPT06, CK06c, CFRR10, DPLW08, DPMP^{+}11a, DPMP^{+}11b, Erl06, FSTS09, GPAS^{+}06, GWK09, GAB02, GCC^{+}02b, JP07a, Kah03, KG04, KPSW10, KHI06, KHS07, LOH05, Lor01a, Lor01b, MZCC09, MKR^{+}04, Omi00, PS09c, RCJ^{+}08, SPSV^{+}04, TPCB12, THX^{+}07, VFC08, ZCCL10, dSDL04, Bau99, BAWH00, Cho02a, Cho02b, GH94a, GH94b, HS95b, HS97, HHW01, JY96, Kim05a, LCHS96, LD05b, LD05a, M^{+}06, PPF^{+}00, SBH98, Slo99, ASE^{+}07, BBK12, BKL^{+}12a, BKL^{+}12b, BAP^{+}05, CGMP11, CFRR10, DPLW08, DPMP^{+}11a, DPMP^{+}11b, FSTS09, GAB02, GCC^{+}02b, GCC^{+}02a, MKR^{+}04, BPS06, LD05b, LD05a, LZ05, MB87, PPF^{+}00]. Networks [ASE^{+}07, ASP01, APPE09a, APPE09b, AH02, ACK00, AHJ07a, AHJ07b, AHJ07c, AH07, Beh03, BPS06, BAP^{+}05, BKM06, BY06, BGW06, CPS05, CSPC09, CRJ11a, CRJ11b, CSJD06, CZ02, CJ06, CK06c, CCBW11, CCK95, CGS07, CMS11, CSRS08, Dad01, DPLW08, DY06a, Del00, DDI10a, DDI10b, DDI10c, Dol09, Dor02, ENO12, Fek08, FLL03, FCLG04, FLR11, GD10, GG98, GCC^{+}02b, HS08b, HJMP06, HDGK11a, HDGK11b, HDLM09, JWH06, KG04, KWW04, KMF06, KT06, KMS06, KSDO06b, KSDO06a, Kou11, KO07a, KPR07, KNK08a, KNK08b, KMN94, KR06, KCK08, LV07a, LWZ09, LPR12, LYCV08, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, ML04, MST12, MSY08, MBH^{+}09, MVCM10, MW11, MA03a, MDY07, MCSY07, MCP93, MS95, MMMD97, MSA99a, MSA99b, MP01a, MP01b, MA03b, MA03c, MKR^{+}04, NKC09, NBK04]. Networks [NW10, NR04, NR06, NLS08, ORB^{+}09, OSC^{+}08, OM98, ODV08, PAD08b, PB09, PH12, PBK03, PvS07, PMS08, RS00, RST09, RR04b, RS09b, RGLA09, SP09, SNČ^{+}12a, SPCG07, Sav06, Sch10a, SSN08, SM06, SE10, SKB10, SS06c, SMCM07, SWB06, STTP02, SD10, SZT^{+}08a, SZT^{+}08b, TSE^{+}09, TMN09, VC03, Ver08, VDÉ^{+}12b, VDÉ^{+}12a, VD12, WW07, WSS04, WLY05a, WLY05b, WLY05c, WYZ06a, WYZ^{+}06b, WYP12a, WYP12b, Yas00, YWG04a, YWG04b, YHZ09a, YHZ09b, YHZ09c, ZOCJ07, ZSM^{+}09, ZLK10a, ZLK10b, dSADM07a, dSADM07b, ABLM05, AW90, BDR99, BC05, BWT95, CHPW05, CJ05, Cho88, CGP03, Cov99, DH05, DBH01, DKOZ05b, DKOZ05a, Fur95, FU96, GGHN97, GSP01, Kay03, MMD05, MSH85, MB01, Mar03, MNW01, MTW05, MBC^{+}97, Omi00, Pla96, Pri91, PPF^{+}00, RMV96, RKM06, SS05b, SLL01, SC05]. networks [Tri98, VH04, ZG03, Zui99, Zwi89, von96, Dad01, MSA99b, Sch98c]. NetworkTechnologies [IHX^{+}08]. Neubiberg [BE87, BEM91]. Neuchatel [CM95]. Neuenahr [Jan88, MT86].

Neuhold [HNR05]. Neumann [KKS11a, KKS11b]. Neural [APPE09a, APPE09b, AMT03, AMNT06, AHJ07a, AHJ07b, AHJ07c, Beh03, CPS05, CSPC09, CRJ11a, CRJ11b, DDI10a, DDI10b, DDI10c, Dor02, ÉEMS04, GG98, HDGK11a, HDGK11b, KSDO06b, KSDO06a, KNK08a, KNK08b, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, MST12, MT02a, MSY08, MMMD97, MSA99a, MSA99b, MP01a, MP01b, MA03b, MA03c, ORB^{+}09, PMS08, SPCG07, Sav06, SM06, SE10, SZT^{+}08a, SZT^{+}08b, VDÉ^{+}12b, VDÉ^{+}12a, WLY05a, WLY05b, WLY05c, WYZ06a, WYZ^{+}06b, WYP12a, WYP12b, WAW01, WPE05, YWG04a, YWG04b, YHZ09a, YHZ09b, YHZ09c, ZLK10a, ZLK10b, dSADM07a, dSADM07b, BWT95, DBH01, DKOZ05b, DKOZ05a, Fur95, FU96, GGHN97, Kay03, MCP93, MS95, Pri91, WS00, von96, AW90, AMT03, AMNT06, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, IDMY08a, IDMY08b, KWCW06a, KWCW06b, KWCW06c, LLC09a, LLC09b, LZK11a, LZK11b]. Neural [LZK11c, MT02a, OM98, PKM04, Roc92, WMB10b, WMB10a]. Neuro [KKC09a, KKC09b]. Neuro-Information [KKC09a, KKC09b]. Neuroergonomics [SEG09]. Neurons [MP01a]. Neuroscience [SEG09, WAW01, ÉEMS04, MMMD97, WPE05]. Neuroscience-Inspired [WAW01]. neuve [UPB07, Qui98, QS00]. Nevada [GHP09, PS98a, Ram01]. NEW2AN [ABK12, BMK08, BMK09, BDK10a, BKH11, KHI06, KHS07]. Newark [AYB95, GPCL10]. Newcastle [CK01, DMRS06, FHT05, GH06, HS08a, KLT^{+}12, KU12, KP11]. NEWCAT [Hau86]. Newfoundland [NH03]. Newport [CC07, LN09]. Newton [Naz94]. Next [ASE^{+}07, ABK12, AH07, BMK09, BDK10a, BKH11, CA08, aCF06, CW07, DPLW08, GVCA07, Kou11, MCA08, OSC^{+}08, PZ92, RST09, RR04b, SSN08, VC03, KS05b, MSY00, BMK08, CHCA09, EKM06, FKK09, HG02, KHI06, KHS07, PT99, RS00, SS91]. next-generation [MSY00, CHCA09]. NF [AC09, NQR09, VS09]. NfC [Bru06b]. NFM [BHHJ11, GP12]. NFPSLAM [PJC^{+}12]. NFPSLAM-SOC [PJC^{+}12]. NGC [CH01, SCKR03]. NGC’99 [RF99a]. NGI [aCF06, GVCA07, KS05b]. NGITS [EKM06, FKK09, HG02]. NGITS’99 [PT99]. NH [MRvM07]. Nha [TS10]. Niagara [BFvdM10, LV02, Ngo09, STY^{+}07, THX^{+}07]. Nice [Aya95, ADS09, ADGM12a, ADGM12b, ADGM12c, BGH06a, BR89, BDH^{+}99, Cea76a, Cea76b, CP07, CDS09, DFGN12, FZ86, Han07, LPZ12, Sd98, YLS^{+}12, YHV12]. Nicosia [ABHT08, Dik04, MEK03, VJ11]. Niigata [SHN07]. NIJ [Che03a]. Nijmegen [BN94, DTMH10, HP95, Vyt91]. ninth [Mad89, NS82]. Nisan [Gol11b]. Nissim [GKKW09]. Nixdorf [MMR93]. Nizhni [Mal03]. Nizhny [HKLP12]. NJ [AH96, BF12, BRT11, BB11, CJRR07, GKL09, GJRR11, GM08b, HMT10c, MS96b, MS11a, ZGK^{+}07, AYB95, AJRS03, Gan96, GF97, IWC04]. NL [YDW08]. NLDB [BIMW12, HRM^{+}10, HMMW10, KSS08, KLM^{+}07, KFMM06, MMM05, MMM11, ABJ02, BKM01, MM04]. NLP [IK12, LRH10, SGPP06, Chr00b]. NLP4DL [BCG^{+}11]. NM [ASAM07, COT98, PN93]. NMA [BDGN07, DDK11, Dim03]. NMELP [DPP97]. Noam [Gol11b]. NoCoDA [CVLL12]. NODe [AMU03, Cha03a, KMTU03]. NOLISP [CHG^{+}07, FZJE^{+}05, FZJE^{+}06, SCZ10, TGAH11]. Non [ACG^{+}07, CS00a, CHG^{+}07, FZJE^{+}05, FZJE^{+}06, Gor02, Rus05, DPP95, DPP97, Küs01, Rei89]. Non-Classical [CS00a, ACG^{+}07]. Non-Linear [CHG^{+}07, FZJE^{+}05, FZJE^{+}06]. Non-monotonic [DPP95, DPP97, Rei89]. non-native [Gor02]. non-standard [Rus05, Küs01]. Nonclassical [PW92b]. nondeterminism [Dob81]. Nonlinear [CHG^{+}07, SCZ10, SFZE07, TGAH11, Naz94, CEFZM05, FZJE^{+}05, FZJE^{+}06]. Nonmonotonic [BS11b, BGLT05, BBS07, DF11, ELS09, KI01, LN03a, DFN97, EFT01, GLP99, MNT95, BJS93, DJS91, Sch97a]. Nonverbal [EFZKM07, EBAH08, EVV^{+}11]. Noordwijk [Ame91b]. Noordwijkerhout [Pet03, PT91a, PT91b, Sch95, dBdRR89, dBdRR91, dBdRR94, vLdB80]. Nordic [AJN12, JMK09, JC12, Lau12, SSB90]. Nordkirchen [FS99b]. NordSec [AJN12, JMK09, JC12, Lau12]. Norfolk [GP12]. normal [Nij80]. Norman [PHZ11]. Normative [GM06a]. Norms [ACÇ^{+}12, BPD06, CvRVSN12, DFPV11, HMBD09, NVSB^{+}07, PAV^{+}10, SPON08]. Norte [CDGW10, CEM^{+}06]. North [Der89, McC97, RZ94, RS97b]. Northern [BH06c, BW10b, BW01, BLS02, MA03a, WR07]. Norway [ÅBE12, AK10, AR04, ABS91, AÖ12, AB03, BB07, BdB08, BK05b, BCH^{+}05, BS06d, Col07, CDW86, DFL^{+}10, Der92, DS11b, Din04, DFD04, DW06, Fom06, GKS^{+}07, GW09, GK90, GN88, GP98, Hal00, HOD96, Hel94, JK04, JP05b, JMK09, Kap10, KDS01, KS04b, KZ97, KH11c, KOS07, LZ08, MØ05a, MT08c, MMK08, MBJ^{+}07, MMH07, Nos88, Ode04, OST11, PSO05, PGBK04, PRH10a, PRH10b, Pie05, PRR05, RS00, RK93, RJS^{+}08, SHJ09, SZR^{+}08, SPH07, SJW^{+}12, SG03, SMMG01, TMLL09, TMS12, THT03, WGD05, Ytr06]. Norwich [Bac00, YWRS11]. NOSSDAV [LG95, SBC^{+}94]. notation [BM99b, BH95, BHT97, BFH98]. Notes [BG12a]. Notre [AKL^{+}99]. Nottingham [AM07, BGHJ07, DMS04, SOS96]. Nova [Deh97, Tom92]. Novel [Jac09d, JP03]. November [AO02, AAW04, ACvLdR03, ACdR^{+}08, AKZ01, ACN^{+}07, ABD^{+}09, AS11a, AST05, ABG09a, AdBB12, A^{+}08, ANS04a, ANS04b, ANS04c, AKU01, ACK^{+}07, ABKT11, ABCWM99, ASP01, ADP09, AF11, ACW07, Alo05, ABB10b, ACJ09, ABD04, ATM09, ABH^{+}10, AB04, AT12, ABGL09, Ari02, ACL^{+}04, AOM07a, AOM07b, ALL^{+}10, BC09a, BBCL08, BST03, BKA11, BCS09, BCC06, BBA12, BMSMS04, BFF^{+}10, BB90, BF08, BPS11, BSW95b, BS11d, BS11e, Bau99, BCE09, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BJM09, Bha95, BR02, Bla90, BC10b, BMO02, BKW09b, BCC07, BM02b, BCK10, BH05, BS01b, BDW97, BLPW02, BkHJ^{+}04, BHLP07, CL08a, CFZ10, CZF10, CNP^{+}08, CSW06, CD09, CC07, CVV08, CBNR02, CC91, Che99a, CPGL05, CC12, CYZ12, CL04, Chi04, CJ06, CKO01, CKM93]. November [CER79, Cod01, CG88, CW89, CUdlC04, CN05, CNmWH^{+}05, CKTW12, CKL05, CS12c, CH01, CDA^{+}06, CGBDA08, CEER96, DGT10c, DK90, DG06b, DSB04, DSS11, DMB^{+}11, DB04, DGJ^{+}05, DE05, DG10, DWWL09, DLW12, DEQ92, DBM09, Din04, DO09, DO11, DW03b, DZ10b, DW07b, DBB^{+}08c, EKS10, EGO09, EEKR00, EP03a, EP03b, EG97, EOR06, ET01b, ÉEMS04, FV01, FS07b, Fei03, FMW01, FNY93, Fon00, FRL07, Fos99, FLMK06, FU96, FMY04, FKLY08, GRT02, GLCT10, GdATM05, GRG06, GM07a, GGS09, GO09, GKL^{+}11, GMBM05, GHI^{+}12, Gol94, GW98, GHP09, GLW88, GHH^{+}06a, GSS00, GP02, GW12, GAR03, GAR04, Gue05, GRR05, Gue01, GP09c, GM10, HRK^{+}07, HOQ97, HY12, HK05a, HKM95, HS00, HV06, Her92b, HP09, HJW00, HK05c, Ho12a, Ho12b, Hol07, Hol08, HM09, HS11, HM04]. November [HM12c, HNPP07, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, HJ00, Hut94, Hut96, HDLM09, ICSY07, IRM93, IFGH10, IFL^{+}12, Ish99b, IDMY08a, IDMY08b, IWC04, IY95, IN11, Jan93, JS01, JKB96, JNS05b, JDL11, JRJ05, JC12, KN09a, Kam99, KJ09, KMC^{+}01, KMS^{+}11, KM96, KKH09, Kim12, KKS11c, KKS11d, KKS11e, KKS11f, KjHGG11, Kob06, KH11a, KH11b, KKC09a, KKC09b, KPR07, KS08b, KJS01, KMS10, KBC05, LV12, LCD^{+}09, Lai03, LOX99, LSS02, LB05, Lau00, LM98b, wLN99, Lee01, LB03, LL03, LhK03, LN09, LHH10, LRG04, LSY^{+}06, LL04, LRL98, LYL06, LH06, LLWL09, LDXZ10a, LDXZ10b, LLTY12, LWRdR11, LGM11, LZK11a, LZK11b, LZK11c, LKK10, MTT98, MT09b, MZ99, MEK03, MOBS06, MEAC04, MTOK06, MT07a, MF08, MSW05, MPvH04, McK99, MNPW07, MTS^{+}03, MT^{+}03, MT05a, MTH05]. November [MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MTH08, MDH09a, MDH09b, MHD09, Mel07, MMG10, MM82, MMR93, MS99b, MMU96, Mit06, MS00c, MSB07, MNTY05, MB97, MBH^{+}11, NNS03, NMM10, NR07, NKN^{+}05, NHJ12a, NHJ12b, Nie09, NRR12, NY93, NK99, NABK12, NQR09, NBM09, NSS03, OB07, Oho03, Oka03, OMMSP06, OPS11, OBM11, PR01, PKM04, PCG^{+}05a, PCH^{+}06, PCN10a, PCN10b, PCN10c, Par02, PSST07, PV00, PSS^{+}10a, PSPH^{+}10a, PSPH^{+}10b, PdRM^{+}12, PTB99, PDMFF12, PW92b, PV02, PH02, PK03, PSM10, Pet02, PSN95, PZ09, PJ97, Pol97, Qin10, QOZ01, QSWL11, Ran93, RVY05, RK06, RL06, RAM05, RF99a, RR04b, RBsC^{+}06, RCLE05, RSvdP05, Rov98, RV98, RMK07, SNY06, SW04, SFS11, SL11, SK11, San02, SRS03]. November [SL05, Sat97, SDG^{+}07, SZ92, Sha07, SBC^{+}94, SJ06, SKSA12, STA^{+}11, SAG10a, SAG10b, SF96, Smo97, SS08b, SLF91, SC96, SDH03, Sub05, SHRM06, SLM07, SHŚ^{+}07, TKSN11, TNI09, TGAH11, TDK^{+}10, TdRM^{+}09, Ued10, UBW06, VM99, VDPD96, VC03, Vet11, Wal01, Wan04, WTZ^{+}04, WmCL07, WT07, WAR94b, WB09, WJP07, WZ11, WMB10b, WMB10a, X^{+}12a, X^{+}12b, XZCL07, XBM12, YKKZ07a, YKKZ07b, YNB08, Yam08, Yan11a, YK82, Yao95, YKF97, YŞ03, Yi05, ZRP^{+}09, ZRP^{+}11a, ZRP^{+}11b, ZL96, ZS07, ZSO99, ZSP^{+}04, ZW09, ZYRH06, ZF05, dBBGdR03, dBBG04, dBBGdR05, dBBGdR06, dBBGdR07, dBBHL10, dO98, dRWK^{+}10, vES10, vSH06, vdHI99, vdHP12, vdL04]. November/December [PTB99]. November2 [MS06a]. Novgorod [HKLP12, Mal03]. Novi [CIT11, Ong11]. Novosibirsk [BBP96, BBZ00, BBZ01, BZ03, CVV12, FMRW09, Mal01, Mal09, PVV10, VV07, WPZA11]. Novy [KN09b, ČGH^{+}11, GKB^{+}08]. NPC [AS11b, CLWC08, DSZ10, JRJ05, LJJG07, PZYS12, JGXC04]. NS [BJ12a]. NSF [FMY04, MDH11a, Oka03, ZGK^{+}07, Che03a, Heb95, KRMM^{+}02]. NSF-ARPA [Heb95]. NSF/NIJ [Che03a]. NSS [XBM12]. NSW [PSNS99, SMS12, PS96b, VPM97, Yao95]. NT [LwCM00]. NTG [Lei74, Lei74]. Number [Bue04, FK02, HMT10b, HHLR08, HPP06, vdPS08, AH94, Bos00a, Buh98, Coh96]. Numeric [HNP12]. numerical [Col84, Cow77, WL03, AFKL04, BDGN07, CKLM09, Dim03, DDK11, KU94, LVW05, MVW09, VWY97, VWY01]. numerics [ZVU99]. Nuremberg [ACdR^{+}08, HK05a, SO07]. Nurnberg [Han81]. NV [BBKP06, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BFJM09, FILP06, GVO^{+}10, GSYZ10, PRK12]. NY [AP10, AN07, ABRS06, BBH^{+}10, BGKY08, BDDS05, BGHS05, BBH08, CR05a, CM05a, DLT05, DGJ^{+}05, DIS11, DCFY10, FRA05a, FRS05, FW07, Fre98, GW11b, HMRT05, HR06, IKY05, KJR05, LOTY05, MMM10, MMO^{+}11a, MMO^{+}11b, MAFS08a, MAFS08b, MA11, MMG^{+}01, MBA05, Ngo09, NKN^{+}05, OGM05, Pac05, Per11a, Per11b, RMR05, Sch01a, SKPG12, UDF06, WPM05, YDKM06, vEHD05, AH94, BN04, Cle92, Dwa00, Heb95, Kap92, LHK02, Mur99, Pin95, SC00, Van02, Zuc03]. NYU [Yao82].

O [Bos96]. OAGM [PPBL12]. Oak [Cow77]. Oakland [SH07]. OAMAS [VASP09]. Oaxaca [LO10]. Oberlech [MB87]. Oberpfaffenhofen [HSS07, SLP05, SFV^{+}11]. Oberwolfach [BFS92]. Óbidos [BP11]. Object [Ame91a, BdB08, BHS07, BM11a, BDG12a, BHCL87, BJ07, BM98, BB02, Car03, CW02, D’H10b, Dro09, Ern07, Fri06, GN88, GW06a, Heu04, Jae01, Lam00, Mad92, Mez11, Mül02, NNS03, Nob12, Ode04, OKL04, PHSZ06, QCDM06, Tho06, Vit08, WL04, ADR01, AM97, Ber00, Bla05, Bre91, BRR97, COT98, CTT93, Coi96, DKM91, Dit88, Gue99, Ish97, Jul98, KDS01, Knu01, LMV95, Mag02, MTO99, Nie93, Olt95, Pap95, SZ05, Vel94b, VT97, dBdRR91, BPR02, BU94, BGY96, BBC04, Ceb08, CNY95, CW02, DB98b, Eug09, Fro02, FM96, GNR94, Heb95, HM02, KLPP03, MMM01, MØ05a, MD99, NY93, NG10, PZH96, SC07, TWN92, TP94, WL04]. Object-Based [BdB08, BJ07, GW06a, NNS03, SZ05, BGY96, CNY95, GNR94, TWN92]. Object-Centered [Lam00]. Object-Orientation [OKL04]. Object-Oriented [Ame91a, BHS07, BM11a, BDG12a, BHCL87, BB02, D’H10b, Dro09, Ern07, GN88, Mad92, Mez11, Mül02, Nob12, Ode04, Tho06, Vit08, WL04, BM98, Car03, ADR01, AM97, Ber00, Bla05, BRR97, COT98, CTT93, Coi96, DKM91, Dit88, Ish97, Jul98, Knu01, LMV95, Mag02, MTO99, Nie93, Olt95, Pap95, dBdRR91, BBC04, Ceb08, Eug09, HM02, KLPP03, SC07, WL04, BU94, Fro02, MMM01, MØ05a, MD99, TP94].

Object-Oriented. [BPR02]. Object-Relational [Jae01]. objective [Nes95, Nes95, Sch01b]. Objects [AdBB12, BdBC01, BBH^{+}10, PAC08, PH02, PD04, PF06, PF08, PF10, PFM12, SST00, dBBGdR03, dBBG04, dBBGdR06, dBBGdR07, dBBGdR08, dBBM09, dBBHL10, AFS89, Atz93, DF01a, ET01b, GJ86, NL01, Rei93, dBBGdR05, AMU03, BV11, DZ10a, Dit01, FN12b, RME04, Vit10, VJ99]. OBPDC [BGY96]. Occasion [AG09, ADM11, ACG^{+}07, AAN09, ACÇ^{+}12, ADR08, BS11b, BV92, BDR10, BDFM12, BKT12, CRS11, CLKK07, CS12b, DDM08, DKN12, DKS12, EMO10, ELS^{+}10, FJM06, GKKW09, HL05, Hei12, HHG11, JLW07a, JL11, JPR04, KK11c, KR11, Len04, LLM09, Mah04, MM09a, MRU12, Nac12, Pal09, PQR11, RSB07, SPG10, SW10, AJS05, BZ83, Der04, HNR05, HS05b, KMPR04, KMO^{+}05, MvOvRdV05]. OCL [CW02]. OCLWS [Bru06b]. OCS [CW11]. OCSC [OZ09, OZ11, Sch07]. Octeville [ELNM08]. October [AP93, ABB12, APR^{+}06, ABS01, ALS^{+}11, ABG^{+}06, AGP08, ABK^{+}09, Agu12, AKT06, ALS^{+}05, AMU03, ASP01, APO96, AV07, AAB^{+}06, AH02, APSZ02, AS11b, AL03a, AFG^{+}11, AMCL02, AYB^{+}09, Ant97, AP11, AM04, Apt00, AGS91, AHW07, AJ94, AWA^{+}11a, AWA^{+}11b, AOSV08, AH07, ACR12, ACHH12, AJN12, AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, ABRS06, ADK04, BSMM04, BV02, BM96, BE11, BBCL08, Bai05, BLS06, BCT02a, BPP91, BL12, BHKR11, B^{+}12a, Bar03, BG09, BBAB^{+}07, BGP08, BCT02b, BL04a, BPR08, BDG12a, BHGP02, BDCM04, BG11a, BFGL10, BKH^{+}09, Bha07, BZU07, BH92, BTBP^{+}08, BTPPS09, BBC^{+}08, BBD^{+}09, BBR^{+}10, BJN03, BL03, Bod96, BMW10, BB06, BTF02, BBR88, BBR89, BBR90, Bor91, Bor92, Bor93, BK96, BH02, BP05, BGL09, BHL11, BM03c, BC11b, BRT11]. October [Bra73, Bra80, BFvdM10, BKNS06, BW05b, BGR08, BRP09, BL05c, BR98, Bru06b, BK04b, B^{+}12b, BD08, BH11, BTMI11, BJ01d, CLKL09, CHKL10, CBGCCZ12, CDP02, CDP^{+}05, CL08a, CK04a, CK04b, CE95, CNX05, CLWC08, CHD07, CS03, CVLL12, CCK^{+}08, CM12, CG02, CGBZ08, Cha03a, CSR08, Cha09, CL10, CRW08, CNA08, CvSW04, CY12, CH09, CBN99, CDY^{+}11, CP10c, CCK95, CER79, CGS07, Coe98, CMS11, Col02, CML^{+}10, CGS02, CLS07a, CMCS11, CLP05, CEKR02, CTS07, CC97b, dRCVT11, CKL05, CMOZM11, CMM01, CDS09, CFS06, Cup00, COB^{+}08, DGSM11a, DGSM11b, DWSK05, DH05, DVVA08, DH01, DSZ^{+}11, DF01b, DFR02, DSM94, DLNU11, DDFM05, DMB^{+}11, DHRT10, DMW00, DPV11, DGM09, DKM^{+}05, DDJ00, DV07b, DSM03, DOS99, DC00, DGR06, DS11b, DdCF98, DPP98, DKB03, DLO03, Dos93]. October [DL81, DA01, DG05b, DFGN12, EHP94, EFM04, Ehr98, ENR83, EO94, Ehr04, ERRS10, EHH12, ERK07, EHM11, EOSW07, ÉEMS04, EP06, ETG02, EBAH08, EV09, EKS00, FWY05, FGH98, FK87, FD03, FWZH06, FSM03, FKB02, FV10, FdCLJ^{+}11, Fic03, Fil91, FD93, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, FTZ08a, FTZ08b, FTZ08c, FTZ08d, FR99b, FKBA12, FT04b, FCLG04, Fre98, FN12a, FKKB^{+}05, FGTZ08, FSP09, FM91, FMC12a, FMC12b, FMC12c, GG05a, GHJ06a, GHJ06b, GHJ06c, GP11, GCB^{+}05, GCJB09, GBH11, GLP12, GE08, GJ86, GPCL10, GLMP07, GJT03, GLZZ09, GKC^{+}12, GPAD08, GT03, GM08a, GM03, GGHN97, GMP06, Gho10, Gie08, GH12a, GHSW04, GOZ05, GL05, GK01, GKW02, GPB02, Gom01, GS95, GM06b, GZ06, GL08, GTY03, GSS11, GN98]. October [Gue04, Hag02, HI93, Ham89, Ham91, HLR^{+}98, HLC^{+}12, HLMW06, HHP09, HU08, HSSY05, HJMP06, HDBM11, HK10a, HK01, Her00, HT05, HN12, HML97, HRT^{+}03, HF97, HSS96, HGZ05, Hof97, HLW08, HKT12, HVT12, HMG03, HY03, HKS09, HHG11, HW05, HSTU99, HST07, HSVZ10, HKM^{+}11, IFL^{+}12, INR07, IK12, IHQ96, Jac04, Jaf99, JMBS07, JL08, Jan89, Jan92, JSZ95, JSdSA05, JFBH08, JP03, JDL11, JHC02, JPXS04a, JPXS04b, JGXC04, JTB^{+}05, JMK09, JC12, JMT11, KSHS01, KS08a, KM11, KCR04, KS03b, KMD02, KKM04, KV04, KP92, Kie12, Kim05b, KSB11, KWCW06a, KWCW06b, KWCW06c, KW90, KSUZ11, KO11, KP01a, KBHSW06, KRMM^{+}02, KHW03, KKS10, KvdM00, KS07a, KFLS07, KBČ^{+}12, KS12c, KMTU03, KH11c, KKDV02, KPR07, KS91, KNP06, Küh07, LLS00, Lag05, LYL11, LZL11]. October [LNS06a, LNS06b, Lau12, LG94, LMW01, LhK03, LYMM05, LSL^{+}09, LMMN10, LWDM12, LM06b, Lei95, Len93, Leo04, LPZ12, LSHB07, LM97, LST^{+}05, LSYO08, LHDD08, LCF^{+}04, LMT00, LM03, Liu01, LMA08, LWYN09, LR09, LC11, LB78, Lom93, LLMR09, LQO04, LH08, LRS98, LKZCS05, MCA08, MB82, MKPV05, MMD05, MP98, MP06, MSG07, MG08, Mal02, MBS11, MPF95, MAD^{+}04, MSS12, MS06a, MS10b, MS10c, MS12b, MSZ90, MdRIR08, MNW01, MF01, MTV09, MPT10, MMR91, MP09, MSAdAS05, MNPW07, MT05a, MTH05, MT06a, MT06b, MTH06a, MTH06b, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, MRSV07, Mel07, MM10, MNW08, MW08, Mil12, MPPU11, MCSY07, MKR07, MTC^{+}08, Mon96, MBF08, MSH05, MSS04, Müh75, MZF94, MZWZ05, MC05, Nad99, NDB94, NYHO07, NZZ11, NdMO03, NT06, NV08]. October [NLKJ06, NKC09, NWHR06, NV01a, NS81, NLS08, NMP^{+}04, NSL09, OiDP98, OAdGBP00, OYY03, OPP04, Pac05, PS04a, PFCS05, PMÓ06, PB07, PT05, PBH03, PT92, Per11c, PDK^{+}10, PCG^{+}10, Pet03, PU04, PRH10a, PRH10b, PHH10, PNB04, PB09, Pfi00, PMDK00, PBK03, PB97, Pol86, PS09b, PS07b, PT91a, PT91b, PT11, Pri06, QQ11, QGZ03, RSV07, RS00, RZ91, RZ94, RS97b, RO00a, RT07, Rau04a, RRR01, Reu01, RS12, Ric02, RRM06, Ric98, RLT^{+}08, RB10, Ros03, RT09, SNY06, ŞSW07, SdSMO12, STC04, SVS02, Sch93a, SvS00, SS05b, SW03b, Sch92, SNZ08, SS09b, SVV12, SM91, SLH05, SHW09, Ser12, ST02, SW08, SA96, SdDK91, SSD^{+}08, SKPG12, SLC01, SCR06, Sie75, SJT05, Sle92, SCH04c, Smi05, Smo97, SG03, SPSV^{+}04, SLLS03, SPC^{+}08, SMK02, SLPM96]. October [Squ09, SXC05, SS06c, SS99, SvdMOP06, SS06d, SWB03, SRM^{+}11, SSKR02, SGD09, Str99, SRL12a, SRL12b, SRL12c, Sur02, SA04, SC05, SSZ12, TM09, TLC06, TT95, TSG00, TvdBI02, TLZ^{+}08, THL11, Tan11, TTP05, TV94, Tha96, Tho11, TG12, TSS11, TLJ06, Tom91, Tor93, TN08, TND10, TSK92, Tou96, TYNM05, UWB^{+}06, VJKT00, Van05, VEG^{+}11, VJAA03, VA95, VU09, VMBMN04, VH04, VHM07, VGB11, Vog88, VSTN06, Vol93, VLY09, WC95, Wal04a, Wan04, Wan05a, WLC^{+}06, WDGL10a, WDGL10b, WGLL10, W^{+}12, WZ09, Wel01, WC98, WVD02, WBD08, Whi00, WCK11, WGL96, WF77, WFB12, WKB04, WLS89, WKS^{+}10, XCHZ11b, XCHZ11a, XBS^{+}10, XCR11, XHXL08, XLHX08, YNB08, Yak00, Yak02, Yak04, YN06, YRWS11, YLS^{+}12, Yas00, YGGÖ05, Yor08, YHV12, YSR^{+}06, YKM06, YGL^{+}10, YLC^{+}10, ZdC08, ZPT^{+}06, ZLJ05, ZZZL12]. October [ZGT05, Zho01, ZRTS03, ZLLC09, ZYH03, ZZTG07, ZZLL10, ZY01, Zil87, Zim92, ZBY07, dBBGdR08, dBBM09, dLWC03, dS02, tTVH^{+}12, vB05, vDLJ11, vF91, vSN01, vdBGG10, vdHXMW08, vdL02]. October13 [SG09]. October13-15 [SG09]. Odaiba [VJAA03]. ODBASE [MT05a, MT06a, MT06b, MT07b, MT07c, MT08a, MT08b, MDH09a, MDH09b, MDH10a, MDH10b, MDH^{+}11b, MDH^{+}11c, MTP^{+}02, MTS^{+}03, MT04c, MT04d]. ODBIS [CCG08, Col07]. Odense [Kuc11, PM96, Vid12, WL93, ZBH12]. ODEs [CKPB07]. ODIS [MHD09]. Oeiras [BC06]. Off [Mor06]. Off-the-Shelf [Mor06]. Offshore [MJ07]. OH [BG11a, BD08, CGBZ08, CU08, XKGS12]. Ohio [HSB^{+}96]. Ohrbeck [ENR83]. OHS [RA00, RTB02]. OHS-6 [RA00]. OHS-7 [RTB02]. OIS [RBsC^{+}06]. Oisterwijk [de 98, dS02]. OK [PHZ11]. Okinawa [Sad13]. Oktober [Bra73, Müh75, Sie75]. Oldenburg [DO02]. Ole [OKL04]. Ole-Johan [OKL04]. Ollon [SF08b]. Olomouc [Bec79]. Olympia [LMS96]. Omaha [BAdVR08]. OMDoc [Koh06b]. On-Demand [BCL03, SF10]. On-line [Pfl02]. ONDM [TNP^{+}07]. ONISW [HRK^{+}07, HP09]. Online [AMZ09, BS09, BJ10, EP06, EK07, JSO04, JSO11, KS08a, SOP12, BJK06, PSO05, FW98, OZ09, OZ11, Sch07]. Ontario [BGS11, BK02, CAD^{+}02, DSS03, DMRS07, GS94, HA00, JSK10, Lau93a, McC96, OK96, RWY97, ST01, TM99, TG04, VY01, WY98, YP01]. Onto [DMB^{+}11]. Onto-CoM [DMB^{+}11]. OnToContent [MTH06a, MTH06b, MTH07a, MTH07b, MTH08, MHD09, MDH10c]. Ontologies [CPSA10, CCG08, DCMS09, GPB02, PPCS12, SPS09, BCBG05, CSCR08, Col07]. Ontologies-Based [Col07]. Ontology [PST11]. OOER’95 [Pap95]. OOIS [BPR02, BB02, KLPP03]. OOOP [BPD06]. oP [Bui87]. Open [ADM11, ADM^{+}12, BdB08, BJ07, GW06a, Koh06b, NNS03, Nar01a, OS09, PvS07, SC00, MB87, SZ05, Zui99, CKD11, CK12a, RA00]. OpenMP [Cha05b, CZG^{+}08, CGKM11, CMMR12, MdSC09, SHM^{+}10, CGKM11, CMMR12, EV01, EdS08, MCdS^{+}08, Vos03]. openness [RTB02]. Operating [HS74, KN83, SBC^{+}94, BGW93, Her92b, LG95, Ran93, SPZ90, BSG84, GSB78, KG74, KN91, MB82]. Operation [CG12, Kis95]. Operational [SEG09, Rus05]. Operations [APS08, ACK00, AH07, BG09, BK04b, CGS07, DKK08, FKB02, HTMC09, KT06, MCA08, MMD05, MNPW07, NSL09, PMÓ06, SW04, SS05b, SS99, SvdMOP06, MMD05]. OPODIS [ARS09, APW06, ALR11, BBT08, Hig05, LMM10, Shv06, TTF07, PH04]. Opportunities [LLO08, SPC^{+}08, SHRM06, YC11]. Optical [TSE^{+}09, DHO08, DO09, DO11, TNP^{+}07]. Optimal [JN01, Dji89, Kno98]. Optimization [AB11, AMZ09, AC09, AJRS03, AGCP^{+}11, AS07c, BS06a, BN04, BJN03, BB10, BDMS08, Cea76a, Cea76b, CT07, CJRR07, Coe11, CR73a, CR73b, CS02b, CvH07, CC09b, CM10, DGB^{+}06, DBB^{+}08b, DXZ07, DHP09, EFG^{+}09, ES10, FW07, GKS^{+}07, GJRR08, GJRR11, GR06a, GW11b, GJRS12, HM12a, HAB02b, JK00, JLV02, JKRR04, JN01, Kno98, Lin12, LPR08, LMT10, MPK04, MMMP12, MBW08, MH11, MvL10, NQR09, ODP^{+}07, PRV11, PT08b, RR04a, SGG06, SSJR10, SPB00, Stü09, SM09, TDWG11, VW07, VAM06, WZD11, WD10a, WD10b, YDW08, vHC08, vHH09, vLS11, AG01, BC95, BJK06, BBRM98, Bre97, CJRT05, CM03b, CAZ05, CBW99, CMQ96, DJRZ06, DJNR09, DBB^{+}04, FP96, FP90, Fon03, GJRT01, HS12b, HJRS99, JR98, JNS05b, JRR03, JK05b, PVV05, PR87, RG05, Wal99b, Waś96, Zit01]. Optimization [GR04, Has96, Mar75, Col84, FS03, TZ89]. Optimizations [PA01]. Optimized [Bra76]. Optimzation [BJP12]. Orange [DF06, MS11a]. Orcas [Bor94]. ORDAL [BM94]. Order [BW03b, BNUW09, CMT02, MMT08, NPW02, SB07, AH00, BDH^{+}99, BJ01b, GN98, GF97, HM05b, JS94, MC94, SS89, SWAF95, SBG04, Vog92, dBdRR89, vGH96, DHMM96, God96, Har79, Hee94]. order-sorted [SS89]. orders [dB93, BM94]. ordinary [Chi79]. ordination [Oss99]. Oregon [Auc98, Ban94, Buh98, Kah87]. Orford [FGK^{+}10]. organic [BL05b, MSUB04]. Organisations [ABD^{+}12, Sch96, Sch98a]. Organising [BSH06, BHJY07, SKRZ04, BSA05]. Organization [BEM91, BE87, LS89, Lom93, Mor89b, SKB98]. Organizations [BPD06, CvRVSN12, DFPV11, HM03, HMBD09, MH04, NVSB^{+}07, PAV^{+}10, SPON08, AFM01, CN98, RB00, Str99]. organized [LdRV94, VASP09]. Organizing [LH11, PM09, GT05, BG11b, HS08c, HK07, KH12, SH09, WMdLA10, dMS06]. Orient [MSWZ06]. Oriental [LMA09, MSWZ06]. Orientation [OKL04, TAS05]. Oriented [Ame91a, BP07, BHS07, BM11a, BDG12a, BHCL87, BB02, D’H10b, DB98b, Dro09, Ern07, GN88, KOFJ09, KMK10, KMSJ11, LCH^{+}12, Mad92, Mez11, Mül02, Nob12, Ode04, RA06b, RA07a, RA07b, RO09, Tho06, Vit08, WL04, WH11, ADR01, AM97, Ber00, Bla05, BM98, Bre91, BRR97, Car03, COT98, CTT93, Coi96, DKM91, Dit88, Gue99, Ish97, Jul98, Knu01, KT10, LMV95, Mag02, MTO99, Nie93, Olt95, Pap95, RA06a, dBdRR91, BCS09, BKM08a, BGHS05, BBC04, Ceb08, DL06b, DGT10c, DPZ12, DR09, Eug09, FL09, GRB^{+}07, GMO03, GHSW04, GGS11, HM02, HKM^{+}07, JSV11, KMMN11, KBHSW06, KHSM^{+}08, KLPP03, KHK^{+}08, KVMH09, KLN07, LLTY12, LP08, LGS09, MWYF10, MRY^{+}11, MZ06a, OWPY03, PZ07a, PJC^{+}12, SC07, WL04, WG11, BCT06, BU94, CW01, Fro02, GOW03]. oriented [HR99b, MMM01, MØ05a, MD99, OGM05, TP94, WWC02]. Oriented. [BPR02]. Origins [HLR00, Sch03c]. Orlando [BGP08, CPvdH03, DG02, Duf11, FR98, FK09, Har11a, Jac11a, Jac11b, Jac11c, Jac11d, JL00, Kur11, LSW00, dVLKR^{+}08, Mar11a, Mar11b, OZ11, PB07, Rau11, Rob11, Rol98, SS11c, SF11, Shu11a, Shu11b, SL99a, SS11d, Ste11e, Ste11f, Ste11d, Ste11c]. Orléan [DLJ12]. Orleans [ACW07, Bro94, DW03a, Ibr95, LPA00, Mai88, Mai90, MMR91, MW03, Raj09, Win09, DLM07]. ORM [MTH05, MTH06a, MTH06b, MTH07a, MTH07b, MTH08, MHD09, MDH10c, MDH11a]. Orsay [GJ93, KJ88, MVN86]. Orthogonal [Ahn04]. Osaka [AHMT08, Asa96, BR96a, CWiM^{+}09, IMK06, Luo12, NK99, WSTI08]. OSC [DHO08, DO09, DO11]. OSF [Sch93a]. Oslo [AÖ12, BB07, BdB08, DS11b, GW09, GN88, GP98, HOD96, JMK09, KH11c, LZ08, MØ05a, MT08c, Nos88, Ode04, OST11, PRH10a, PRH10b, RJS^{+}08, SHJ09, SZR^{+}08]. Osnabrück [HBE07]. Other [AD06, CHCA09, DT06a, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, JvdAB08, JvdA09, JBK09, JDK10, JDK12, JvdAM^{+}12, JDAW12, KY07b, MMO^{+}11a, MMO^{+}11b, NBGA08, OA07, BRR87a]. OTM [MT05a, MTH05, MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MTH08, MDH09a, MDH09b, MHD09, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, MTS^{+}03, MT^{+}03, MTC04]. OTMA [MDH10c, MDH11a]. Ottawa [AMW07, BG07a, DSS89, DFK91, DSS91, DSS03, EW03, FK10, JSK10, KR06, MDY07, MZ04, OYA04, OK96, SM01, Tan11]. Ottmann [KSW03]. Otzenhausen [Ric98]. Oulu [AMS06, BO00b, BKS05, OKHH^{+}08, RKM^{+}12, RYG11]. Ouro [TDWG11]. Outcome [HBH08]. Outdoor [DFP^{+}12]. output [Bre97]. Outreach [CC12]. Outsourced [MJ07]. Over-constrained [JFM96]. Overlays [SHS10]. Overview [CKV08, ABGP05]. overviews [dBRdR86]. Oviedo [Cue03]. Oxford [BCG02, BMW93, CG12, CMM09b, Che11, DG07a, DP12, GW96, Gib12, GH04, HM05b, JPD04, JJ03, LS06a, ML95, Ong05, PHJ03b, PHJ03a, RSL01, Sie86, Tah09, TMN09, YI91, dMGFS11]. OZ [Van05].

P [Die04, Tha84, ST87]. P-functions [Tha84, ST87]. P.R [DZ94]. P2P [HKW11a, HHMT12, LMT^{+}04, MDH10c]. P2P-CDVE [MDH10c]. PA [ABF^{+}07, AKM10a, AKM10b, ABB^{+}10a, BL05a, BdQ11, BT06, CBT07, GH07a, JMP12, KS06, KR12, LPH09, LS06c, MBG12, MBP^{+}11, MGH09, RZ12, SW09a, SS09d, vHH09, APSZ02, AL03a, AP04a, BCT02b, BRW85, Bro92, BCR03, DDJ00, FGH98, GMV03, KT96, KC84, LS95, LM96, LK00, MNW01, McA00, Mos01, O’H98, Pal00a, San02, VKJ03]. PAASC [MRY^{+}11]. PACCF [YCC^{+}08]. Pachuca [SAG10a, SAG10b]. Pacific [ANS04a, ANS04b, ANS04c, AIL12, AH07, BK05a, BHH08, CL06a, CWZ^{+}11a, CWYC12, CLY02, CCH02, CYCL09, CChL^{+}10, CC12, CPC07, CKH10, DSZ04, DLW^{+}07, DFW^{+}11, FG96, GHR89, GGS09, HCL05, HK05c, HZ08, Ho12a, Ho12b, HhKRT06, HTMC09, HXC^{+}08, HCS11a, HCS11b, IAL^{+}07, Ish99b, IS02, JE06, KT06, KL02, LM98b, LB03, LCHS08, LXH^{+}12, LS09, MCA08, MJR04, MR07a, MS00b, MWK^{+}09, NcC99, NKLC06, OS03a, QLK^{+}10, QLK^{+}11, RK09, RK12, SWJX12, SS06b, SLC01, SXC05, TCHB12b, TCHB12a, TKCH09, WHL09, WSTI08, WKK98, XCR11, YW06, YZC^{+}07, YKF97, YX04, YLLZ04, YY01, ZYRP10a, ZYRP10b, ZS00, ZGY04, ZTY^{+}05, ZYBX08, ZO10, ZLS^{+}06, ZLY07b, ZYRH06, CWL01, McK99, TLC00, Wha03, ZZ99, Zho01, ZZO03]. Pacific-Asia [CLY02, DSZ04, HCL05, HCS11a, HCS11b, TCHB12b, TCHB12a, TKCH09, WSTI08, WKK98, ZYRP10a, ZYRP10b, ZLY07b, CWL01, TLC00, Wha03, ZZ99]. Pacific-Rim [LXH^{+}12]. Pacifico [IZ98]. package [Cha85]. PACS [FV04, FV05, FV01, FV03]. PaCT [Mal05, Mal07, Mal09, Mal11, Mal95, Mal03, Mal97b, Mal99, Mal01]. PaCT-95 [Mal95]. PaCT-97 [Mal97b]. PaCT-99 [Mal99]. Paderborn [BR92, Bil97, Bun96, DHMM96, EEKR00, FLR11, KP92, MHA09, MMR93, MM96, MC89, MF02, MS08b, MYS10]. PADL [CP10b, GS09b, Han07, Hen06, HC05, HW08b, RL11b, RZ12, DW03a, Gup99, Jay04, KR02c, PC00, Ram01]. PADO [DF01a]. Padova [Apo93, AM04, BDCM04, Cou93b, SA04]. Padua [AFP^{+}10]. PAIGOS [Ull08]. Pairing [GP08a, JMO10, SW09b, TOOO07, GP08a, JMO10, SW09b, TOOO07]. Pairing-Based [GP08a, JMO10, SW09b, TOOO07]. PAIS [CWC^{+}07, CLZ^{+}09]. PAISI [CWZ^{+}11a, CWYC12, CYCL09, CChL^{+}10, YZC^{+}07, YCC^{+}08]. Paisley [FTC^{+}10]. PAKDD [BHH^{+}06, CHB^{+}12, CWiM^{+}09, HCL05, HCS11a, HCS11b, LYT06, NKLC06, TCHB12b, TCHB12a, TKCH09, TNA^{+}10, WZH^{+}07, WSTI08, ZYRP10a, ZYRP10b, ZLY07b, CLY02, CWL01, DSZ04, TLC00, Wha03, WKK98, ZSD03, ZZ99]. PAKDD-98 [WKK98]. PAKDD-99 [ZZ99]. PAKM [KR04a, RK06, Yam08, KR02a]. Palaiseau [BP09b, Nie09]. Palermo [AGS91, BU94, DR05, DPP09, DFT12, GM11a, KGG^{+}11, MMMO96, PS11a, RL11a]. Palm [DG05b, Wil99]. Palma [BR97, BL05c, CGBDA08, EP06, FSdlPT11a, FSdlPT11b, LS04b, LSF07, Luo04, Luo05, NL01, PH02, PD04, PSS98, TJ08, VY93]. Palmas [ANF08, CAS^{+}09b, CH99, DPA05, DPA07, LMB08, MÁ05a, MÁ05b, MDP04, MDBF01, MDPQA09, MDPQA12a, MDPQA12b, PMD90, PMD94, PMD97, TGAH11, VSH11, vdL98, vdL01]. Palo [GW98, NW97, SW09b, SC96, ZZ04a, ZG03, dKIM^{+}07]. PAM [CU08, Dov05, KP10b, MTU09, SR11, TR12, UPB07, BP04a].

Pamplona [vdHS10]. Pan [KMN94]. Pan-European [KMN94]. panacea [AM03]. panels [Fro02, MMM01, MD99]. Panhellenic [BH05, MEK03]. Panorama [SPSV^{+}04]. Paper [Sad13, AKR03]. Papers [ACG^{+}07, ADB^{+}12a, ADB^{+}12b, Ant97, BBP^{+}12a, BBP^{+}12b, BHU03, CUDL02, CR04a, CR01b, CDNS00, CKOS09, HO03, IDMY08a, IDMY08b, Jón12a, Jón12b, JS02, KKS11a, KKS11b, KKS11c, KKS11d, KKS11e, KKS11f, KH11a, KH11b, KMSS02, KKC09a, KKC09b, Kut12a, Kut12b, LRS03, LM00, LSDN04a, LSDN04b, MW08, MDPQA12a, MDPQA12b, PAC08, PSPH^{+}10a, PSPH^{+}10b, SPG10, WDKW10a, WDKW10b, WDKW12a, WDKW12b, Ytr06, ZiTM10a, ZiTM10b, ZiTM10c, BW91, Cho02a, Cho02b, DN93, GOW03, JRR03, MDBF01, PDH99, PDH01, PDHS03, PMD90, PMD92, PMD94, PMDA96, PMD97, PMDK00, ST88, WWC02, ARS03, AKK04, AS11a, ABG^{+}06, AYB95, Ada96, ABH12, AZ96, AGP08, AS99, AM03, AdBB12, ABF^{+}07, AKU01, AK03, AKT06, ACK^{+}07, ABKT11, AMU03, ABMS06, AV07, All96, ACW07, ABdJS10, AKR^{+}12, Alp11, AFKL04]. Papers [ANR99, AM07, AC09, Ama08b, APW06, ABJ02, AS98, AGT98, AT07, APN08, AMNT06, Apt00, AFR^{+}04, AS10, AÖ12, AS12, ABGL09, AK12, AS07b, Ari02, AL11b, AOSV08, APV09, AM02, AAR05, AJ01a, AMM99, AHSC11, ACHH12, AJN12, AKS09, ABRS06, ABFR07, BBMB^{+}08, BBD^{+}10, BW03a, BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, BdBC01, BS09, BJ10, BFGM05, BYLJ11, BL12, BN94, BKLM07, BHKR11, BHJKPG06, Bar03, BGH06a, BF08, BDE12, BB91, BRS^{+}07, BUG93, BPR05, BM11a, BDG12a, BBDS12, BHGP02, BS06b, BDG^{+}12b, BP09a, BC96a, BCD04, BDd09, BHM^{+}04, BHMS07, BMS04a, BCM00, BCV06, Ber03, BB01b, BLSS11, BM99b, BC93, BVAT12, BY07, BZU07, BMS95, BGS11, BBZ01, BBH01, BS10, Bla03b, BJN03, BPT06, BRST08, BK02, BB10, BGU94, BGU95, BDC12, Bod03]. Papers [BRU06a, BFSN10, BPD06, BCC^{+}12, BMN05, BKW09b, BJ01a, BLvdH10, BB06, BFMS08, BFJM09, BDDS05, BDDS06, BDDS07, Bor93, BGM94, BG07a, BCC07, BGL09, BGJ11, Bos00b, BR07a, BFT^{+}08, BDN08, BRM06, BKL^{+}12c, BKM01, BGKS04, BDGN07, BR07b, Bra07, BC11c, BKM07, BBT10, BB11, BGHS05, BBK^{+}11, Bre03, BAdVR08, BHLvdP07, Bri95, BGY96, BEFP08, BZ03, BKM06, BKM08b, BtHRBB12, BSH06, BHJY07, Bru06b, BW09, BS12, Bru04, BS01c, BBB06, Bun96, BKB03, BR96b, BC98, BE01, BD03, BR07c, BY06, BTMI11, BDLMM12, BHM^{+}02, BFOY04, BH06d, BCDN07, BJ01d, BGH06b, But10, BGW06, CS00a, CJ11, CG12, CLMP02, CCS04, CC95, CKP^{+}12, CCJS07, CKD11, CL11, CK12a, CRL01, CV04a, Can12, CGL^{+}09, CBG^{+}10, CBS^{+}12, CHB^{+}12, CV08, CMSM^{+}10, CP12, CS03, CS05a]. Papers [CPSA10, CGMP11, CW94a, CM95, CA08, CR01a, CAS^{+}09b, CW11, CY03, CMZ99, CM03a, Cha03a, Cha09, CWiM^{+}09, CH04, CPZ07, CLZ^{+}09, CF09a, CY10, CY11a, CYZ12, CHG^{+}07, CD00b, CM09a, CHZ08, CBS^{+}11, CW94b, CEFZM05, CKO01, CK06c, CS10, CFG96, CBN99, Chr01, CCMR02, CCMR03, CCMR05, CCMR06, CCMR07, CCMR09, CCMR10, CMMR11, CCMS11, CMSA12, CSN14, CCG08, CSY09, CY11b, CW01, CGP03, CFVV^{+}12, CFVCF^{+}12, CHD98, CV10, CVV12, Coe11, CPvdH03, CF09b, Col07, Col02, CML^{+}10, CDN12, CUdlC04, CLS07a, CM00, CMCS11, CFP^{+}09, CM09c, Cor02, CKTW12, CF10, CS12c, CvRVSN12, CMOZM11, CDM04, CRW^{+}12, CGBDA08, CLBP11, CEER96, CKLM09, CE00, DLFP02, DFL^{+}10, DR02a, DKY06, Dam97, DYCD09, DGT10c, DM06, DG06a, DDS12, Dan12]. Papers [DF10, DGT96, DDE04, DSLT08, DSRW08, DSLT10, DSHL11, DS09a, DM01b, DSS11, DDHP05, DPC^{+}07, De 07a, De 10a, De 10b, DFPV11, DW12, DS09b, DHtM^{+}12, DGG^{+}11a, DGM09, DV09, DG10, DEG11, DFP^{+}12, DBB12a, DLW12, Des09, DJS06, DJNR06, DLN10, DGSN11, DKN^{+}11, DEM96, DR06, DR09, DP10a, DGWB09, DPA07, Die02, DD07, Die03, Dig04, DvEF06, DGR06, DGDM^{+}11, DMRS06, DMRS07, Dim03, DDK11, DS11b, DS03, Din03, DG09b, DKN12, Dit01, DPP95, DPP97, DPP98, DFN10, DJ08, DB07, Dol09, DO11, DS11c, DFT09, DP97, DMW06, DHMM96, DW06, DW07b, DKR02, DKN02, Dun09, DS06, Dwa00, DS07, EJM10, EDGP11, EL07, ELS12, EKF12, EKO08, EO94, EEKR00, EKP^{+}07, ET01b, EPVW10, EG10, EP06, Erl06, EK07, ENO12, Esc10, EIRW06]. Papers [EFZKM07, EBAH08, EV09, EHMM09, ECV^{+}10, EVV^{+}11, EEM^{+}11, Est95, EH12, FRS08, FBKS03, FBSMS08, FV01, FV03, FPFR05, FPRAW04, FS07a, FZJE^{+}05, FZJE^{+}06, Fei03, FR01a, FRS02, FRS03, Fek08, FKK09, FGB11, FLVS11, FMZ05, FL09, Fia99, FS07c, FKM09, FdCLJ^{+}11, FdCLN^{+}11, FPCK11, FST09, Fle99, Fog94, Fog95, Fog96, Fom06, Fon00, FLUV02, FS07d, FS08, FS09b, FS10, FL11, FKM08, Fre98, FN12a, FPRS06, Fuc10, FLMK06, FLT07, FKLT08, FCDB07, Fur95, FU96, FMY04, GVO^{+}10, Gae96, GKL09, GML^{+}08a, GPCL10, GANACBR10, GANACL11, GANACBdCdV12, GAL12, GCFA12, GVCA07, GY08, GS01, GLV09, GMV03, GKS^{+}07, GMP06, GW03, GKT09, GKT10, GKST11, GKS12, GG02a, GHP^{+}11, GPR^{+}12, GGS09, Gho10, GCK06, Gie08, GKL^{+}11, GH05, GH12a, GH06, GPM98, GMO03]. Papers [GHSW04, GLM^{+}08, GT04, GOZ05, GGS11, GSS09, GGHS07, GSLM12, GM99, GK02, GR06b, GP08b, GCC^{+}02a, GW06b, GNTF02, GO06, GH93, GKMN10, GHI^{+}06, GVT^{+}11, GAR03, GAR04, GS06, GB07, GHP^{+}08, GM07b, GDG11, HM11b, Hag02, HENR03, HDC99, Han09, HLR^{+}98, HLC^{+}12, Har01, HK12, HKM95, HKI^{+}09, HNRW06, HOD96, Hee94, HDBM11, HM03, HBN06, HHLR08, Hic04, HBH08, Hig05, Hil06, HRT^{+}03, HRR^{+}06, HPS09, HR08, HCF06, Hof97, HhKRT06, HCR06, HŽ07, HNQ08, HI10, HPRS06, HM05a, HMMW10, HNP12, HZB07, HM11c, HNPP07, HGBF12, HGGG12, HGPH12, HMBD09, HKS09, HHG11, HDLM09, HMSU04, HdMKS12, HKM^{+}11, IM10, IB07, IPSK04, IMW04, IS11a, IS11b, IST07, IGN05, IFV07, IHX^{+}08, IWC04, IT03, IKKU08, IS12c, JRSN09, JMBS07, JSS08, JSO04, JSO11, JP07a, JLST06]. Papers [JLM00, Joh03, Joh01, JP11a, JJF03, JDMB08, JDMB10, Jou11, Jue04, JP12b, JY12, JRR03, KEDW07, KS03a, Kah03, KG04, KS08a, KN09a, KKP02, KCR04, KA12, KUA00, Kap08, KT96, KS09b, KW07a, KM08b, Ker12, KS12b, KL12, Kie12, KYL07, KKH09, KSB11, Kim12, KM04, Kin08, KDS01, KHG01, Klu97, KTB09, KK11d, KBHSW06, KHSM^{+}08, KRMM^{+}02, KC00, KW10, KS07a, KS07b, KBČ^{+}12, KMS12, KLT^{+}07, Kow01, KMTU03, KH11c, KMVY06, KMYJ08, Kuc̆02, KBC^{+}09, Küh07, KRS02, KZ10a, Kur10, KCO04, KŽ10b, KCK08, LSJ06, LJS08, Lac10, LV12, LYL11, Lam96, LSV06, LVS08, LS10a, LSW01, LSW02, LSW03, LO11, LN04a, LS12a, Lau01, Lau12, Lau93b, LS04a, LLP99, LT11a, wLN99, LL03, LhK03, LYY07, LC09, LH10, LY12, LMSS07, LOST04, LT05, LOTY05, Lei95, Leu08]. Papers [LS05, LM08a, LWM^{+}08, LLLL08, Leu03, LLP^{+}07, LOS05, LLM^{+}08, LCF^{+}04, LL04, LM96, LAF^{+}10, LMP04, LMCH12, LYCV08, Lio04, LMWY04, LMW08, LMW10, LMW12, LN01b, LCY03, LLL06, LL06, LLO08, LW12, LK04, LK07, LK09b, LOH05, Lop06, LH08, LS06b, LSU07, LMŠT01, LP08, LGS09, LSW08, LS09, LV07b, LCY^{+}11, MWD06, Mac12, MMMP12, MP06, MSG07, MG08, MBS11, MBW08, MEK03, MY06, MMBND07, MSRP98, MS06a, MWY01, MVW09, MŠKL02, MOBS06, MT02a, Mar11c, MRU12, MF01, MVMP^{+}04, MVOF08, MS05, MP10b, MR12a, MS00a, MTV09, MPT10, Mat02, MZ04, MS03b, MPR07, MRY^{+}11, MRPM10, MRP11, MPR12, MMM10, McK99, MFG12, MLTC11, MMG10, MCSN10, Mey95, MS99a, MT02b, MJ07, MNW08, MB09, MSH08, MMG^{+}01, MCSY07, MV12a, MA05c, MNLM06, MP12, MOT12, MK01, MTC^{+}08]. Papers [MSB07, MvL10, Moo01, MS11a, MF06, MG07b, MDP04, MDBF01, MDPQA09, MBS05b, MK03b, MBJ^{+}07, MK12, MS02c, Mug97, MOTN07, MTNL12, MPS04, MZ06a, MGSM12, MNTY05, MJL02, NMM10, NP09, NP11a, NP12a, NZZ11, Nar01b, NZS^{+}06, NSS^{+}07, Ngu06, NT98, Nie09, NR06, NLS08, NS99, NVSB^{+}07, NG10, NvRSS11, Nür03, ND04, NH03, Nyb08, O’H98, OROD07, O’S03b, OO12, ONPR07, OFR09, OLL10, Oka03, OS07, OBS12, OW09, Ölv10, OTZ01, OPT01, OPP04, OSV11, OBM11, OK96, OSRS07, PT95, PSP^{+}02, PAV^{+}10, PZ07a, PM08, PHZ11, PP01, PVW10, PR07, PT09, PJC^{+}12, PAD^{+}08a, PDML11, PPCS12, PH04, Par98, PMMR06, PS12a, PH10, Pau03, PPJRN^{+}10, PYLW08, PW09, PA03a, PvES12, PP12b, PRT06, PS07a, Pet01, PBGK02, PBGK03, PCG^{+}05b, PGG^{+}06, PCG^{+}07]. Papers [PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, PK03, Pet03, PNPP12, PU04, PTZ03, Pet02, PNB04, PKSS02, PVV10, Pod95, Pol01, PBRB08, PT06a, Pri03b, PQ05, Pro11, Pue07, PS09c, QLXL07, QCDM06, RPR08, RM09, RM97a, RS99a, RLL04, RUL04, RRU05, RTZ08, Ras01, Rau04a, RKM^{+}12, RTA95, RTB02, RBF06, Ret97, RSS06, RN11, Ria08, Ria09, RtTMP10, RtTMP11, RtTM12, RK09, RGW01, RBTD04, RL11a, RSL01, Rob06, RH01, RHD08, RB10, RF12, RBSL06, RM04b, RME04, SST00, SS08a, SNY06, ŞSW07, SM11a, SSVW12, SFS11, SHN07, SÇS12, San02, SH12a, SINK08, SGG06, Sav06, SMB03, Sch10a, ST08, ST11, SKP07, Sch03a, SKP10, SC11, SJDT08, SNRG11, SSN08, SNZ08, ST12b, SHW09, SG09, SCL^{+}06, SLL^{+}07, SYY^{+}08, SPÖ^{+}10, She03, SKPG12, SBD03, SA06, SPON08]. Papers [Sil01, SM11b, Sio10, SCD^{+}10, SA12, SHP^{+}05, Smi05, Smi06, ST07, SSM08, SCZ10, SOP12, Som96, SMF^{+}11, SKK04, Squ09, SKCK07, SS04, SP07, STWS07, SG07, SBF08, SZ11, Stü09, SV01, SM09, SHŚ^{+}07, SSZ12, TvdBI02, TT10, TBC07, TNA^{+}10, TKSN11, TPCB12, Thi10, TGH06, TP09, TC98, TT06b, TST09, THR03, Tsu08, TtHVS06, TWSN07, TNGK08, UWB^{+}06, UBW06, VS09, VAR12, VMvN12, VEG^{+}11, VY01, VFC08, VU09, Vet11, VHM07, Vid12, VSMS12, VV07, VS05, VASP09, VKG12, VWY01, WS02, WK08, WR01, Wan98, WmCL07, WJL^{+}12, WZH^{+}07, WSTI07, WW03, Wei97, WAR94b, WvdH03, WPM06, WPM07, WBD08, WMdLA10, WG11, Wii04, Wii05b, Wil99, WL03, Win04, WPH03, WKB04, WHR10, WPZA11, WK06, WY98, WVDS05, WCGB05, WKD08, WDPW02, WDPW04, WDM06]. Papers [WDKW08, XW11, YNB08, Yal10, Yao95, YKF97, YLWY06, YJKK06, YJKSW10, YC11, YSL12, YMY^{+}10, YY09, YKMJ09, YAKMJ10, YAKMJ11, YAKM^{+}12, YP01, YLL09, ZSD03, ZSSM03, ZL97, ZSM^{+}09, ZCDT10, ZZZL12, ZLY^{+}07a, ZY01, ZHP12, dCdF^{+}08, dBBHL10, dGEK11, dGN12, dKIM^{+}07, dMGFS11, dS02, dSOSR04, dSOSR06, tCZ07, tHBP08, vB97, vDKS11, vES10, vEDA04, vKS12, vLS11, vdBK05, vdBGG10, vdHHsHD06, vdHCD07, vdHS10, vdHIP11, vdHP12, vdL02, vdL04, vdMS11, AJS05, AMW07, A^{+}08, ABK05, ADB^{+}05, AL12, AGL06, Bai05, BDGM05, BK05a, BBH^{+}05, BB05a, BR06, BK05b, Bir07, BC05, BL06c, BT06, BTF05, BS06d, CP06a, CHPW05, CR05a, aCF06, CZ05, Cha05b, CSV05, CKMS05, CKL05, CRYS09, DS05a, DLT05, DS05b, DPA05, Dik04, DL04b, DRS05, DOSY05, DLW06, Dur12]. papers [ELM05, EW05, Eta05, FV04, FV05, FRA05a, FLS06, FRS05, FFRS05, FMO05, FPMW06, FNY93, FKKB^{+}05, Fri04, FLMS05, GS05, GN05, GHMT05, GT05, GF12b, GAN05, GM05, Gue05, GRR05, HS12b, HH04, HN06, HSSY05, HPR05, HRTR05, HW06b, HNW04, JSdSA05, JNS05b, JY14, KIRK05, Kim05b, Koh06a, KS05a, KS05b, Kra05, KBC05, Lag05, LBCO04, Li96, LVW05, LBO05, LY05, LMT^{+}04, LMW06, LA05, Mac06, MML97, MRK05, MNS05, Mar05a, MPPJ^{+}05, MTW05, MSS05a, MBA05, MSS05b, MZWZ05, NOOS05, OGM05, OFK05, Pac05, PTRH05, PC05, PY05a, Per05, PSO05, PGBK04, Pie05, PT06b, RMR05, RWY97, ŞS05a, SVV12, SHGB05, SDH05, SLB05, SXC04, SKY06, SS06d, STLK05, TLC06, TBG05, TMP04, TYNM05, TAS05, VZ06, Van05, VT97, WPM05, Wii05a, Wri03, vEHD05]. Papers. [RFBD09]. Papers/IICS [BHU03]. Paphos [Apt00, AL11b, ADP11, BJ07, Cod01, EM10, Eug09, Gor10, Gup10, HK12, IR07, KMMN11, MS06a, MP09, MBF08, MV07, Ölv10, Ong10, PJC^{+}12, Pet02, RT10, SEO^{+}09, VAR12, Vit08, Wal01, ZBRL12]. PAPM [DG01, HS02]. PAPM-PROBMIV [DG01, HS02]. Par’96 [BFMR96a, BFMR96b]. Par’97 [LGG97]. Par’98 [PR98b]. Par’99 [ABD^{+}99]. PARA [DMW06, Jón12a, Jón12b, Kåg98, KEDW07, DW94, DMW96, FHJ^{+}02, SMMG01, Waś96]. Paradigm [LLP^{+}07, HK98a]. Paradigms [AM91a, BFGM05, KKW11, KKW12, KPRS07, SS04, SG01, WBHR08, CLT05, Gre94, HG96, SMMG01]. Paragon [She85].

Parallel [AGP08, ADB^{+}12a, ADB^{+}12b, ACW07, AS11b, Ama08b, ABRS06, BEM91, BCRT92, BFMR96a, BFMR96b, BFT^{+}08, BZ08a, BV94, Bur90, CLWC08, CHD07, CAS^{+}09b, CMCS11, DGT10a, DGT10b, DSM94, Die03, DSZ10, DKP00, DLO03, DMW06, DGJ09, EV96, FHJ^{+}02, FR00, FRS03, FS07d, FS08, FS09b, FS10, GPCL10, GVT^{+}11, Han81, HHJ^{+}86, HYPY10a, HYPY10b, HM11c, HC09, IH90, Jae01, JNR11a, JNR11b, JGXC04, JRPP07, Jón12a, Jón12b, KEDW07, KPS12, KKW11, KKW12, KBP07, KKDV02, KKD04, LKD08, LMSS07, LJJG07, LAF^{+}10, LMB08, MTWD06, MCdS^{+}08, MO06, NWL06, ORS89a, ORS89b, PA01, PZYS12, Rau04a, Rol98, Rol99, RWD09, SEL09, TYZ11, WLS89, XCHZ11b, XCHZ11a, XSA^{+}12a, XSA^{+}12b, XZCL07, ZH00, ZJXC03, dBNT87a, dBNT87b, AvLR91a, AvLR91b, AD98, ABD^{+}99, BL92, Ban92, Ban93, Ban94, Bar85]. parallel [BE87, Bee89, Bil97, BRW93, BLKW01, Bos96, Bre97, BGY96, BDW97, CNX05, COT98, CF00a, Cas01, Cha05b, CPC^{+}99, CD01, CM05b, DLV04, DKD05, DMW96, DP97, DLM99, EV01, ES92, FL00, FR95a, FR96, FR97, FR98, FR01a, FRS02, FRS05, FFRS05, FP96, Fer98, Hal94, HAM95, Has96, HGZ05, HSB^{+}96, Ish97, IY95, JRJ05, Kåg98, KBH03, LGG97, Li98, MTO99, MMG^{+}01, MF02, Mue01, PDH99, PDH01, PS97a, PS98a, PD96, Pin95, PR98b, PT06b, RF99b, Sak01, Seh97, Sip97, SL99a, SMMG01, The89, TU94, TV87, Vos03, Wal82, Waś96, ZVU99, AJM87, Ame91b, BG91b, BD93, Bod96, BS94, Bos96, BCRT92, BV94, CYGL04, Che92, CPGL05, CCD^{+}12a, CCD^{+}12b, CFP94, DSM94, DMWW88, DW94, EV96, EBSS98, yF75, FR95b, Fer96, GY03]. Parallel [GYD^{+}06, HI93, Ish94, IH90, IHQ96, LSS^{+}04, Mal95, Mal97b, Mal99, Mal01, Mal03, Mal05, Mal07, Mal09, Mal11, MAB^{+}02, MMR93, Nak92, PCG^{+}05a, Rol98, Rol99, Rol00, RJL^{+}08, RBB^{+}06, SCKR10a, SCKR10b, SDR^{+}00, SM91, STY^{+}07, Vol93, WDPW02, WDPW04, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, YD98, YBL^{+}04, Zim92, ZVU99, vZvdV89]. Parallelism [CNY95, EdS08, MdSC09, SHM^{+}10, JY96, Zil87, Zip92, BEM91, Boc95, Rao95]. parallelization [Bre97, FW92]. parameter [JM78]. Parameterized [BL06b, CF09a, DFD04, GN08, MR12a, RS10, TW12]. Paraty [AJRB09, VGB11]. Parcella [WLS89]. ParDMCom [M^{+}06, THX^{+}07]. Paris [APFV09, AK90, AV99, AAA^{+}11, ABCWM99, AG98, AC09, AB06, ACD^{+}10, Baa07, Ban98, BG91b, BM11a, BV92, BCI85, BHCL87, BMYZ91, BMYZ95, BDN08, BGKS04, BBB06, BDLMM12, CHD07, CS03, CP10a, CL07, Che99a, CHG^{+}07, CCLR98, CG88, Coh92, Coh94, CGM95, CMS11, CBF01, Cou01, Cou05, CBCGA12, DS05a, DJ90, DKLT05, DGG^{+}10, DTF98, ES92, FM84, FW09, Fri01, GPAS^{+}06, GAL12, GL10a, GNR07, Gau93, GM12, GH05, HS12b, HIS05, Heu98, Heu00, Hor00, HM98, KLM^{+}07, KW98, KNP01, KS07a, KSGT10, KR04b, LV12, LSW01, LLP^{+}07, LS89, MT07a, MKZB04, MMK98, MSR99, NP02, Nad99, NNS03, NPPDG06, ND09, NV08, NP10, Omi00, PMA^{+}07, PT08b, Pit91, PMS08, PPF^{+}00, Rob74, Rob80, Roc07, RBC93, SDR^{+}00, SCG98, SDH03, SP12]. Paris [Van97a, Vau98, WZ11, YBM87, Zam08]. Paris-Rocquencourt [APFV09]. Park [BKA11, BR12b, CH98, CMB^{+}08, DJ08, KT96, MTNL12, Pal00a, PT06b, SYNC11, SBG04, SBM11, TDS94, YGE12]. PARLE [AvLR91a, AvLR91b, ES92, Hal94, dBNT87a, dBNT87b, BRW93]. PARLE’89 [ORS89a, ORS89b]. Parlog [Che92]. PaRMA [GHI^{+}06]. ParNum’99 [ZVU99]. parsing [Hau86, Nij80]. PART [BF10, AL11a, AI06, ADG^{+}08b, AHS11a, AHS11b, ANS04a, ANS04b, ANS04c, AMSM^{+}09b, AKM10a, AKM10b, ADB^{+}12a, ADB^{+}12b, ADJ^{+}01a, ADJ^{+}01b, AvAS06, AvASD06a, AvASD06b, AvASD06c, APPE09a, APPE09b, ANS^{+}09a, ANS^{+}09b, Ama08a, AGS^{+}11b, AGS^{+}11a, AHJ07a, AHJ07b, AHJ07c, AAH^{+}10b, AAH^{+}10a, AWA^{+}11a, AWA^{+}11b, AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, Ayk07a, Ayk07b, Ayk09, BBGS10a, BBGS10b, BBGS10c, BPAB07a, BPAB07b, BH06b, Bar11, BS11d, BS11e, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BBP^{+}12a, BBP^{+}12b, BDIR07a, BDIR07b, BKL^{+}12a, BKL^{+}12b, Bir12, BTBP^{+}10a, BTBP^{+}10b, BTCW10a, BTCW10b, BHQ10a, BHQ10b, BvADS08c, BvADS08a, BvADS08b, BPSW06a, BPSW06b, BGMST09a, BGMST09b, CSPC09, CRJ11a, CRJ11b, CK04a, CK04b, CK06b, CK06a, CK10a, CK10b]. Part [CK12b, CK12c, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CPFD^{+}03a, CPFD^{+}03b, CFZ10, CZF10, CDCW09, Cas09a, CCD^{+}06a, CCD^{+}06b, CW09, CCP10a, CCP10b, CCD^{+}12a, CCD^{+}12b, CRS10, CKW11a, CKW11b, CSA^{+}12a, CSA^{+}12b, CMPW12a, CMPW12b, DGT10a, DGT10b, DGSM11a, DGSM11b, DGM08a, DGM08b, Dai07, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, De 07b, DPB^{+}04b, DG12, DMLW11a, DMLW11b, DMLW11c, DCC^{+}10, DBD^{+}10a, DCC^{+}11, DBD^{+}11, DDI10a, DDI10b, DDI10c, DLŠ11a, DLŠ11b, DPMP^{+}11a, DPMP^{+}11b, Dro08, Duf07, Duf09, Duf11, DL07, EP03a, EP03b, EM10, FSdlPT11a, FSdlPT11b, FI08, FMP11a, FMP11b, FMP11c, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, FDC12a, FDC12b, FK12a, FTZ08a, FTZ08b, FTZ08c, FTZ08d, FMC12a, FMC12b, FMC12c, GHJ06a, GHJ06b, GHJ06c]. Part [GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, GTM10a, GTM10b, Gel11b, Gel11a, Gel12a, Gel12b, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, GG07a, GG07b, GG07c, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, GO11, GLC^{+}11a, GLC^{+}11b, Gor10, GGK^{+}09a, GGK^{+}09b, GH07b, GHMV11a, GHMV11b, GHMV11c, Gup10, HLSZ11a, HLSZ11b, Har07, Har09, Har11a, Hen08, HP06, Ho12a, Ho12b, Hof11, HDGK11a, HDGK11b, HYPY10a, HYPY10b, HLI06a, HLI06b, HLI06c, HCS11a, HCS11b, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, IDMY08a, IDMY08b, IS12a, IS12b, Jac07c, Jac07d, Jac07b, Jac07a, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, JNR11a, JNR11b, JNHJ10a, JNHJ10b, JNH11a, JNH11b, JLS11a, JLS11b, JCC^{+}12, JNPV10a, JNPV10b, JNPV10c, JWbG^{+}06, JWG^{+}06, Jón12a, Jón12b, KC11a, KC11b]. Part [KKS11a, KKS11b, KvETvdH10b, Kar09, KHJ05d, KKS11c, KKS11d, KKS11e, KKS11f, KWCW06a, KWCW06b, KWCW06c, KILW10a, KILW10b, Kno11, KH11a, KH11b, KSDO06b, KSDO06a, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KKC09a, KKC09b, KP09, KO07b, KP15, KNK08a, KNK08b, KKGT03c, Kur09, Kur11, Kut12a, Kut12b, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, LNS06a, LNS06b, LTL^{+}11a, LTL^{+}11b, gLPZ^{+}12a, gLPZ^{+}12b, LBP06a, LBP06b, LBP06c, LBP06d, LLC09a, LLC09b, LSDN04a, LSDN04b, LFJI10a, LFJI10b, LJS^{+}10, LSTZ12a, LSTZ12b, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LDXZ10a, LDXZ10b, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, LD05b, LD05a, LHJ08a, LHJ08b, LHJ08c, LZK11a, LZK11b, LZK11c, MF11a, MF11b, Mar11a, Mar11b, MSS12, MS10b, MS10c, MS12b, MdlBP05a, MdlBP05b, MBMS07a, MBMS07b, MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MDH09a, MDH09b]. Part [MDH10a, MDH10b, MDH^{+}11b, MDH^{+}11c, MMO^{+}11a, MMO^{+}11b, MAFS08a, MAFS08b, MKZK10b, MKPZ12a, MKPZ12b, MSA99a, MSA99b, MA03b, MA03c, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, MDPQA12a, MDPQA12b, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, MZ06b, NNS06a, NNS06b, NHJ04a, NHJ04b, NLŚ10a, NLŚ10b, NKJ11a, NKJ11b, NHJ12a, NHJ12b, O’B12, ORB^{+}09, Ong10, OF15, OZ09, OZ11, PM04a, PM04b, PM04c, PM04d, PPCS12, PCN10a, PCN10b, PCN10c, PCN12a, PCN12b, PCN12c, PSDS11a, PSDS11b, PSPH^{+}10a, PSPH^{+}10b, PSD^{+}07, PSM^{+}07, PRH10a, PRH10b, QLK^{+}10, QLK^{+}11, RR08, RDPMA^{+}11a, RDPMA^{+}11b, RR98b, RR98c, Rob11, RCS10, RT10, RST^{+}10b, RST^{+}10a, RKS^{+}12a, RKS^{+}12b, SS11c, Sas05, SCKR10a, SCKR10b, Sch04b, SR07, SEG09, SF11, Sch07, Sei07, Sei12, Ses06]. Part [SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, SvADS07a, SvADS07b, SvADS07c, SvADS07d, Shu09, Shu11a, Shu11b, SAG10a, SAG10b, SYP^{+}05, SS07c, SS07b, SS09c, SS11d, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, SRL12a, SRL12b, SRL12c, SZT^{+}08a, SZT^{+}08b, SvASD05a, SvASD05b, SvASD05c, TST10a, TST10b, TSCW11a, TSCW11b, TCHB12b, TCHB12a, TSJ12a, TSJ12b, TKCW11a, TKCW11b, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, TT10, VRHJ09a, VRHJ09b, Ver00a, Ver00b, VDÉ^{+}12b, VDÉ^{+}12a, WLY05a, WLY05b, WLY05c, WYZ06a, WYZ^{+}06b, WDGL10a, WDGL10b, WYP12a, WYP12b, WMB10b, WMB10a, WD10a, WD10b, WDKW10a, WDKW10b, WDKW12a, WDKW12b, XCHZ11b, XCHZ11a, XSA^{+}12a, XSA^{+}12b, XHXL08, XLHX08, YKKZ07a, YKKZ07b, YHR^{+}09a, YHR^{+}09b, YHZ09a, YHZ09b, YHZ09c, YKU11a, YKU11b, ZYRP10a, ZYRP10b, ZRP^{+}11a, ZRP^{+}11b, ZiTM10a, ZiTM10b, ZiTM10c, ZLK10a, ZLK10b, dA09]. Part [dBM10a, dBM10b, dLZ12, dMS09, dSADM07a, dSADM07b, ADG^{+}08a, AMSM^{+}09a, Bod05, BRR87a, BRR87b, Cer05, Cle99, Deg03, Due04, DG05b, Est98, Fin99, GH03, Gor03, GS00b, HZ05, Han98, Hed03, HM01, Hor02, Hus01, Jäh99, JP04a, KS02c, Kos98, Le 02, Mai00, MY01, NE02, Niv98, NvRSS11, Pez03, Rau11, Sag05, SS09a, San01, Shu07, Smo00, Ste98, Swi99, Tho99, Tiu00, Wal04b, Wat00, WK02, WMS04, Wil01a, AvLR91a, AvLR91b, BHH04b, BHH04c, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BC96b, BC96c, Cea76a, Cea76b, Cho02a, Cho02b, CR73a, CR73b, DPB^{+}04a, DK02a, DK02b, Ekl94a, Ekl94b, GL74a, GL74b, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, KS02a, KS02b, Lam00, Lor01a, Lor01b, MdPA98, MP01a, MP01b, NHJ04c, ORS89b, PT91a, PT91b, STDH02c, STDH02a, STDH02b, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c]. Part [SAB^{+}03d, YWG04a, YWG04b, dBNT87a, dBNT87b, dPMA98]. Part-Whole [Lam00]. Partenkirchen [EB04, SP93]. Partial [DLdS04, Vog92, Zwi89, dBdRR89, DGT96, God96, HMT99]. Partial-order [God96]. Partially [ST12b]. Participation [MT09a, TMG10, TMdB11, TMS12]. Party [Lin03]. Pasadena [BHHJ11, CS11, HW09, PAV^{+}10]. Pascal [QCDM06, Har77, JW74, JW75]. Paso [GLP99]. Passau [BVNW87, Bra96, GLIC10, GF12a, LGG97, LKK10, MW91, MS96a, STP^{+}10, dMB05, dMS06]. Passing [CHD07, CDND11, DKP00, DLO03, KGRD10, KKDV02, KKD04, LKD08, MTWD06, RWD09, TBD12, AD98, BDW97, CD01, DKD05, DLM99, JM78, Koy92]. Passive [BP04a, CU08, Dov05, KP10b, MTU09, SR11, TR12, UPB07]. PATAT [BT06, BR07c, BE01, BD03]. PATAT’97 [BC98]. paths [Akm87]. Pathumthani [CJ06, KL95]. PATMOS [AGCP^{+}11, AS07c, MvL10, PVV05, SM09, VAM06, vLS11, CM03b, HAB02b, MPK04, SPB00]. Patras [Chr00b, EBAH08, PW09, Ria09, SMK06]. Pattern [AMGJ12, AP10, AT02b, BCE09, BC10b, BKLS11, CAD^{+}02, CBBS09, EV07, FL08, FCD^{+}04, GM11a, GHI^{+}12, HWW^{+}10, JFT11, KS12a, KU09, LV06, dVLKR^{+}08, MZ07a, MST12, MTOK06, MTCOK10, MBS05b, NJ09, NJA^{+}11, PH11b, PI05, Per07b, Per09b, Per11b, Per12b, PMS08, RFZ03, RBGS04, RMK07, RSK08, SMD04, SK11, SRS03, SM06, SE10, SK12, SSAP05a, SSAP05b, YKF^{+}06, YZCW07, ZJL06, Ami98, AL01, Apo92, Apo93, AH97a, ACP05, BYCC03, BV05b, Can89, CG94, FC98, FIAP01, FZJ01, GU95, GS00a, HP99b, HV03, HM96, PC99, PH97, PpWR96, PP99, Per01, PR03b, RVY05, STC04, SL05, SMK01, TEB09, AMPT09, COMTLB12, CMM^{+}09a, CNA08, DNS09, DTMH10, FMNS06, GDP07, GRK^{+}10, HSJ07, HAB02a, KSG^{+}09, Kit88, KSH05, KMKP11, LV02]. Pattern [LWRdR11, MdlBP05a, MdlBP05b, MBMS07a, MBMS07b, MTCOBH11, MF11c, MK03a, PBB05, PCPS03, PPBL12, RF01, RWA06, RSV07, Rig08, SKSA12, VSH11, vG02]. Patterns [BV11, FN12b, GE08, JP09a, KKH07, PW03, RDPMA^{+}11a, RDPMA^{+}11b, Sch03c, ÜÇA10, Vit10, ZSSM03, CK93, GP05, HS95c, Ska01, SL99b, SDP97, NJA^{+}11]. Paul [DHS10]. Paulo [BC10b, CG02, SM11b, Sim92, dLWC03]. Paun [KK11c]. Pavia [BSW95a]. paving [Zui99]. Pawlak [PSD^{+}07, PSM^{+}07]. PCI [BH05, MEK03]. PCM [ANS04a, ANS04b, ANS04c, HK05c, HK05b, HXC^{+}08, IAL^{+}07, LXH^{+}12, MWK^{+}09, QLK^{+}10, QLK^{+}11, ZYRH06, CCH02, SLC01]. PCRCW [BS94, YD98]. PDCAT [LSS^{+}04]. PDCE [M^{+}06]. PDE [KSW05]. PDK [RB91]. PDMC [BHLvdP07]. PE [CGMP11]. PE-CRN [CGMP11]. Pedagogically [Ull08]. Peer [AKK04, BDG^{+}12b, DJS06, GCC^{+}02a, Ham08, HT09, HMT10a, HT11, JDMB08, JDMB10, MK03b, MBJ^{+}07, SW05, JM02, MSS05a, MBA05, NOOS05, CR05a, DKR02, KS03a, TAS05, VS05]. Peer-to-Peer [AKK04, BDG^{+}12b, DJS06, GCC^{+}02a, Ham08, HT09, HMT10a, HT11, JDMB08, JDMB10, MK03b, MBJ^{+}07, SW05, JM02, MSS05a, MBA05, NOOS05, DKR02, KS03a, VS05, CR05a, TAS05]. Peirce [Luk97]. Penang [HS00, LLL06]. Penh [HV06]. Peniscola [DR81]. Pennsylvania [Dei90, Gib89, HF97, Tom91]. People [AGB^{+}11, BG09, GKL09, LFK^{+}02, MKZ02, MKZB04, MKZK06, MKZK08, MKZK10a, MKZK10b, MKPZ12a, MKPZ12b, CCK95]. Perceiving [KvETvdH10b, KvETvdH10a]. Perception [ADM^{+}08, Fer08, IS12a, IS12b, SG07, SBF08, Tu99, VEG^{+}11, Lau93b, SK97, SZ00, Tan93, ADM^{+}06, ADM^{+}08, KMMP12, Rob93]. perception-action [SK97, SZ00]. perceptual [Rob93]. Pereslavl [Mal07, MT89]. Pereslavl-Zalessky [Mal07, MT89]. PERFORM [HHG11]. Performability [SRT06]. Performance [ABBC10, APBP07, BB95, BH07, BPS06, BAP^{+}05, Bra09a, BKZ05, CT02, CDE^{+}07, CNmWH^{+}05, DDHP05, DPC^{+}07, DKS^{+}07, DN93, DSS^{+}07, GK06, GCC^{+}02b, HS02, HT06b, KS09a, KS03c, KGS08, Kou11, MDY07, MCSN10, MKR^{+}04, MPV01, NP12a, NMP^{+}04, PAD^{+}08a, PDML11, PP12a, PFD^{+}10, PCS^{+}07, PP03, RPBP06, SPBP08, SEO^{+}09, Ste07a, SDK^{+}08, Ste09a, Ste11a, Ste11b, TJ08, Tho11, THX^{+}07, VJKT00, VJAA03, Wol07, ZCDT10, ZJS^{+}02, BPSP06, BKS99, Bee89, BP04b, BHK01, BAWH00, CS87, DG01, ELM05, Fie02, GHR89, HK94, HBS00, HR99a, MR97a, M^{+}06, PTRH05, PDHS03, Pol97, Pol99b, PSS98, PPF^{+}00, RMV96, SKS02, VKV00, YRMD05, CT02, CG04, DN93, DRSS01, HLR00, HHG11, KGS08, LJS08, NP09, NP11a, GH94a, GH94b, HS95b, HS97, HHW01, LCHS96]. performance [SBH98, Slo99]. PerMIn [KMMP12]. permutation [But91]. Pernambuco [BCSW10, CSW06]. Perpignan [CM11, ECB06, TND10]. Persistent [KDS01]. Person [KJR05, LLT^{+}04, BCB97, BS01a, KN03, LST^{+}05]. Personal [CGGO03, COB06, NdG04]. personalities [TP97]. personality [TP97]. Personalization [DKC10, HMPZ09, KCMO11, MMDN12, MA05c, BKN07]. Personalized [FN12a]. Persons [GKL09, ZBW94]. Perspective [AFGvE10, AFGvE11, Bra09b, DP10b, KD11, KSW03, KLLP12, LDC^{+}09, AZ96, GH92, Bra09b, BM09b, DS01b, KBEP03]. Perspectives [HP09, HKV10, JSV11, KM11, LL06, LK04, Mit05, Mit06, MS08a, SGG06, SC09a, WZA10, vdBK05, CS03, dBdRR94, ALS^{+}05, BBP96, BBZ00, BBZ01, BZ03, CVV12, FST01, GV08, KSFN08, PVV10, VV07, BL94, II00]. PERSUASIVE [BR12a, OKHH^{+}08, PHOK10, IdKM06, dKIM^{+}07, BR12a, IdKM06, OKHH^{+}08, PHOK10, dKIM^{+}07]. PerSys [MTH06a, MTH06b, MTH07a, MTH07b, MTH08]. Perth [GF03, PWZ98, Sat97, WR11]. Peru [FB12]. Perugia [De 95, GML^{+}08b, GML^{+}08c, Lee11, Lio04]. PERVASIVE [FSN06, LLT07, AP09, BCC^{+}10, CL07, CM06, CPPB06, FKS10, GGK06a, HMSU04, IPRO08, KLT^{+}12, LCB12, LHH11, MBH^{+}09, MN02b, MSUB04, RKM^{+}12, RYG11, STP^{+}10, TBF^{+}09, WYX08, BL05b, GWS05, HU05, SUW02, FM04, FSN06, FKS10, GGK06a, GWS05, IPRO08, KLT^{+}12, LLT07, LHH11, OYM07, TBF^{+}09]. Peso [CBF09]. PET [BG07a, DM06, DG06a, MS05, DS03, Din03]. Petascale [CGKM11, LMSS07]. Peter [Pal09]. Petersburg [ABK12, BMK08, BMK09, BDK10a, BKH11, BGU94, BKLS11, GSP01, GPS03, GLS05, GKS05, GZSC07, GHH06b, KS10a, KS12c, KHI06, KHS07, KV11, Mal95, Mal99, Mar05a, NM94, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, Vau06, Vor92a, Vor92b, Vor93]. Petri [AB97, BR96a, BRR87a, BRR87b, CD05, CK01, CR04b, DD95, DS98, DRR04, DT06a, EJPR01, ERRW03, EL02b, FW09, HP12, Jen92, JvdAB08, JvdA09, JBK09, JDK10, JDK12, JvdAM^{+}12, KY07b, KP11, LP10a, Mar93, NS00, RR98b, RR98c, RRG85, Roz86, Roz87, Roz88, Roz90, Roz91, RRP91, Roz92, Roz93, Val94, Vog92, vHV08, vdAB03, ADR01, BDR99, BRR87a, BRR87b, CYR02, DT06a, ERRW03, KD99, KY07b, Tau89]. PETS [AH10, BG08, FHH11, FHW12, GA09]. PFE [vdL02, vdL04]. PfHSN [CZ02]. PHD [LMT^{+}04, CLLD09, GHI^{+}06]. PhDS [MTH05]. Philadelphia [ABB^{+}10a, AL03a, AP04a, BdQ11, CBT07, GMV03, GH07a, JMP12, KR12, LS95, LM96, MBG12, MNW01, NSS^{+}07, RZ12, SW09a, San02]. Philippines [HO98]. Phnom [HV06]. Phoenix [BC10a, CTT93]. Photogrammetric [SRM^{+}11]. Phuket [BLJ08, FS07b, YT99]. physical [Nay95, YK82]. Physics [DKN12, DMW96]. Physiological [MT09b]. PIA [SRM^{+}11]. Piceno [BMUV10]. pictorial [Bla80, CF80]. Pierre [CLKK07, GP89]. Piest’any [PR01]. Pillars [ADR08]. Pilsen [HM11a, MM07, MM09b]. PIM [GHI^{+}06, LMT^{+}04]. PinKDD [BFMS08, BFJM09]. Pins [AF02, BP04a, BTBP^{+}08, DC00]. pipelined [RR89]. Piran [KŽ10b]. Piriapolis [BBPP09]. Pisa [AJNJ12, AZ96, APW06, AGM03, Avg12, BBB07b, BEGP04, BEG04, CT03, CM09c, DLV04, DFM04, DEG11, DFG^{+}12, DGLT12, EEK^{+}87, EKLM87, FL08, FP08, FYM12, FT94, GS05, GCC^{+}02b, GCC^{+}02a, GSS11, HKI08, KMVY06, Lev98, MP98, MP10b, MWZ10, MS96c, MO05b, PGM98, Pat02, PdRM^{+}12, PFD^{+}10, PT98, PT97, RF99a, SÇS12, SYMA09, Seb03, SWMM06, TBC07, TV87]. Piscataway [BRT11, FC98, IWC04]. PIT [ADM^{+}06, ADM^{+}08]. Pittsburgh [ABF^{+}07, AKM10a, AKM10b, BCT02b, BRW85, Bro92, BT06, Dei90, DDJ00, Gib89, KS06, KC84, LS06c, LK00, McA00, MBP^{+}11, Mos01, O’H98, SS09d, Str99, Tom91, VKJ03, vHH09]. Pixelization [LLP^{+}07]. PKAW [HhKRT06, KR10a, RK09, RK12]. PKC [CFGN11, Cra08, FBM12, IZ98, JT09, NP10, OW07, Vau05, BDZ04, Des02, IZ00, Kim01a, NP02]. PKC’99 [IZ99]. PKDD [BBGS10a, BBGS10b, BBGS10c, BGMST09a, BGMST09b, DGM08a, DGM08b, DGDM^{+}11, FDC12a, FDC12b, FSS06a, GHMV11a, GHMV11b, GHMV11c, JTB^{+}05, KKdM^{+}07a, RTZ08, BEGP04, EMT02b, KZ97, LGTB03, ZKZ00, ZQ98, dRS01]. PKDD’99 [ŻR99]. PKI [AL06, CZ05, LSF07, MMK08, KGL04]. PL [BSW12, CJE82]. PL-Grid [BSW12]. PL/CV2 [CJE82]. plan [BHGP02, Bee02]. plan-based [BHGP02, Bee02]. planar [Rei93]. Planning [Har11b, Hel08, Her91, Hof03a, Kim05b, Nar01a, Sch03b, SF10, Ull08, BF00, Nar01b, SA97, Vel94a, dPS95]. Plans [Bee00, Had94, von92]. Platforms [Gue01, Jac07d, KS07b, SC00, Ish99b]. Plattsburgh [CP98a]. Playa [WB09]. Playing [CKK^{+}09, Kai11]. Plenary [LABM^{+}12, ZYW08]. Plenary/Invited [LABM^{+}12, ZYW08]. Pleso [GC81]. PLILP [BW92, BP93, GHK97, HP94, HS95a, KS96, MW91]. PLILP’98 [PGM98]. Plzen [MMOS99]. Pnueli [MP10a]. pocket [Hel91]. PODDP [KM04, MNW98].

Podebrady [SS06c]. PODP’96 [NW97]. Pohang [Sch11]. point [Koj91]. Points [CPRS01, KMSS02, Vel94b]. Poitiers [ABB12, ADL05, BV05b]. Poland [Abr07, ALS^{+}11, Abr01, ACJT07, ABT04, BPW06, BDIR07a, BDIR07b, Bih03, BSS11, BMN05, BKW09b, BTCW10a, BTCW10b, BTCW12, BS93, BKL^{+}12c, BKW06, BDW97, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BvADS08c, BvADS08a, BvADS08b, BKB03, CC97a, CKW11a, CKW11b, Deg03, DR02b, DKOZ05b, DKOZ05a, DKN02, FHFL06, Fra05b, GH03, Gor03, Gór06, Hed03, JS05, JNHJ10a, JNHJ10b, JNH11a, JNH11b, Kar77, KY07b, KKdM^{+}07a, KKdM^{+}07b, KY11, KMS89, KPRS07, KRSR11, KPW99, KCK08, KCG09, LMV98, MT04a, MW97, MNW08, MS08a, MHW12, MS11b, NGHJ07, NBGA08, NKC09, OT08, PT95, PS96a, Pez03, PK12, PS98b, RFBW12, RM96, RS99b, RTZ08, RSTZ04, RTZŻ06, RTZZ08, RST^{+}10b, RST^{+}10a, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, SCKR10a, SCKR10b, Ska01, Sko85, SKN05, SKR^{+}10, SSZ12, Tar91]. Poland [TT06a, VU09, Vet11, VLY09, WSGA06, Win78, WDPW02, WDPW04, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, ZCCL10, Dem80]. Policies [Hug03, SLL01]. Policy [KZ10a, DG96, SLL01]. Political [BBGR03]. polygon [The89]. Polyhedral [Heu04, Akm87]. Polynomial [Die04]. polynomials [Sud95]. Pont [Gal97b]. Pont-a-Mousson [Gal97b]. Ponta [CCD^{+}09, FLMG10, MZF94]. Ponte [BO00a]. Pontevedra [BR07a]. Porabka [KMS89]. Porabka-Kozubnik [KMS89]. Poreba [KPW99]. Pori [ACJ08]. Port [DPP98, Fre98, PF06, PF08, PF10, PFM12]. Portability [Cow77]. PorTAL [RM02, Bus85, Bus88]. Portland [AH00, AO02, AS99, Auc98, BR03, Ban94, Buh98, CFB11, EH95, FV05, GJN10, HSB^{+}10, Kah87, KL07a, KR02c, WvdH03]. Porto [ABG^{+}06, BR02, BB06, BJ01c, CKS04, CK04a, CK04b, DN07, DP97, FD93, FCLG04, FLS05, GCB^{+}05, GCJB09, GBH11, GLZZ09, GPP09, JTB^{+}05, MR12c, PDH99, PDH01, PDHS03, PBK^{+}10, PF05, PH06, SS08a, VS09, dO98, dSADM07a, dSADM07b]. Portugal [AS11a, ABG^{+}06, AM03, APO96, AL04, AW90, ARC^{+}07, AP11, AMPT09, ABFR07, BSMM04, BO00a, BSvRW09, BPP91, BA99, BBF^{+}06, BL12, BDPS02, BP11, BDG^{+}12b, BCD05, Ber10, BC06, BMO02, BPS06, BMM07, BB06, BB00a, Bra07, BJ01c, CIM^{+}05, CV07, CCD^{+}09, CKS04, CK04a, CK04b, CK06b, CK06a, CK08, CK10a, CK10b, CK12b, CK12c, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CBF09, CVTP12, CCdS03, CP10c, CC97b, CM05b, DN07, DD97, DH01, DS09a, De 07b, De 10b, DS98, DGWB09, DLGJ10, DP97, DL07, Est98, FBSMS08, FLVS11, FLMG10, Fia99, FV09, Fil91, FD93, Fon03, FCD^{+}04, FLL03, FCLG04, GCB^{+}05, GCJB09, GBH11, GLZZ09, GPAD08, Gen09, GK09, GH07b, Gue99, Han98, HDBM11, HPS09, Hof01, HMBD09, IZP09, JSdSA05, JJF03, JTB^{+}05, KOL^{+}04, KK11e]. Portugal [Kos98, KHK^{+}08, KO07b, KMR12, LO02, LSV06, LVS08, LLPH09, LT05, LP10a, LN04b, LLMR09, LGS09, LV07b, MOP11, MPF95, MBT^{+}03, MSS07, MdlBP05a, MdlBP05b, MM89, MZCC09, MT07b, MT07c, MTH07a, MTH07b, MDH09a, MDH09b, MHD09, MD99, MR12c, MZF94, NSM07, Niv98, Oli00, OMMSP06, PPP07, PDH99, PDH01, PDHS03, PF05, Pet01, PH06, PA03b, QBA01, RM09, RCB^{+}04, RM02, RFZ03, RtTMP11, SS08a, SdSMO12, Sei07, SO09, SKB10, SA12, Smi01, Ste98, SM09, SHO99, TdLdOQ08, VS09, VFC08, VA95, VASP09, Wel01, dSADM07a, dSADM07b]. Portugual [Coe98]. Portuguese [AP11, BPP91, BA99, BCD05, BJ01c, CVTP12, CC97b, FD93, LLMR09, MPF95, MBT^{+}03, MM89, NSM07, PBK^{+}10, PA03b, TdLdOQ08, V^{+}06]. Porzio [JS08]. POS [KDS01]. POS-9 [KDS01]. Posay [Gue90]. Position [Sch03a]. POST [DG12, Fre98, BD08, Gab13, Sen10, TNN^{+}01, WSN^{+}06, Yan11a]. Post-Conference [Fre98]. Post-Proceedings [TNN^{+}01, WSN^{+}06]. Post-Quantum [BD08, Sen10, Yan11a, Gab13]. Posters [DB98b, HM02, MTH05, MTH06a, MTH06b, MTH07a, MTH07b, MTH08, MHD09, MDH10c, MDH11a, Fro02, MMM01, MD99]. Potential [NRW04, BFJV97]. Potsdam [ABMM07, BJ01a, DPW04, ELS09, HLW04, HR08, LS04a, PK00, SHGB05]. Poultney [DG97]. Póvoa [AMPT09, CK06b, CK06a, CK08, CK10a, CK10b]. Power [AGCP^{+}11, AS07c, CM03b, HAB02b, LJS08, MPK04, MvL10, SPB00, SM09, VAM06, vLS11, PVV05, FV01, FV03, FV04, FV05, Mei89]. Power-aware [FV01, FV03, FV04, FV05]. Poznan [Abr07, ALS^{+}11, Kar77, LMV98, MNW08, Sal83, VU09, Vet11, WDM06, MHW12, VLY09]. Poznan-Kornik [Kar77]. PPAM [WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, WDPW02, WDPW04]. PPCP [Bor94]. PPDA [CLLD09]. PPDP’99 [Nad99]. PPK [YDKM06]. PPN [MTH07a, MTH07b]. PPSN [CCD^{+}12a, CCD^{+}12b, RJL^{+}08, RBB^{+}06, SCKR10a, SCKR10b, YBL^{+}04, DSM94, EV96, EBSS98, MAB^{+}02, SDR^{+}00, SM91]. PPSWR [ABMS06, FS05, BHM03, OS04]. PQCRYPTO [Gab13, BD08, Sen10, Yan11a]. Practical [CZG^{+}08, CGG^{+}07, CKOS09, EEM^{+}11, FBHW00, GO96, GLZ^{+}03, GCL^{+}06, LGR^{+}04, MN12b, PvES12, vES10, BJM^{+}05, Bol99, CRW97, CGLR05, Dam99b, Gab97, KHZ82, CP10b, DW03a, GO96, GS09b, Gup99, Han07, Hen06, HC05, HW08b, Jay04, KR02a, KR04a, KR02c, PC00, Ram01, RK06, RL11b, RZ12, Yam08]. Practice [AL11a, AM91b, AM91a, AI06, ABMS06, Ama08a, ASLM09, AZ11, APP12, AL06, BDZ04, BLW09, BW11, BH06b, Bar11, Ben06, Bes07, BFG^{+}12, Bir12, Bod05, BF10, BEGP04, BHM03, CV11, Cas09a, CFGN11, Cer05, ČGH^{+}11, CW09, CDLZ06, CMS08a, Cle99, Coh10, Cra08, DW07a, De 07b, Dec00, Deg03, DG12, DBPZ05, Der04, DLW12, Des02, DO89a, DO89b, Dro08, Due04, DL07, Ehr85a, Ehr85b, EEK^{+}87, EKLM87, EM10, Est98, FI08, Fin99, FBM12, FK12a, FSS06a, GE08, GH03, GAL12, GKB^{+}08, Gen09, GO11, Gor03, Gor10, GBMS11, GP02, GH07b, Gup10, HZ05, Han98, Hed03, HKP10, Hen08, HP06, HHLR08, HJW00, Hof11, HM01, Hor02, HdL12, Hus01, IZ98, IZ99, IZ00, Jäh99, JT09, JP04a, JTB^{+}05, KS02c, Kim01a, KjHGG11, Kno11, KKdM^{+}07a, Kos98]. Practice [KP09, KO07b, KDWW10, LGTB03, Le 02, Lee11, LSF07, MP98, Mai00, MSS11, Mar11a, Mar11b, MY01, MBH^{+}09, MMK08, MZ06b, NP02, NP10, NE02, NKM^{+}09, Niv98, O’B12, OS04, OW07, Ong10, PR01, Pai09, PRK12, Pez03, QBY^{+}12, RWSS12, RR08, RSB07, RBsC^{+}06, RT10, Ros03, Rov98, RSW12b, Sag05, San01, Sas05, Sch04b, Sei07, Sei12, Ses06, Smo00, Ste98, Stu08, Swi99, Tho99, Tiu00, TG10, TZS00, VGP08, VPBŠ03, Van02, Vau05, Wal01, Wal04b, Wat00, WK02, WMS04, WTP^{+}06, Wil01a, WKS^{+}10, YYI^{+}09, YDKM06, ZRP^{+}09, ZBRL12, ZT04, dA09, dLZ12, dMS09, vLIvdH^{+}07, vLMP^{+}10, BSW95b, BD97b, BMS95, BGR03, Bor94, CS03, CFP94, DKN02, FS05, FBKS03, Fre96, GJ93, HMT99, HVT12, IY95, Jaf99, JKB96, KK95, wLN99, Mil12]. practice [MR95, MNS95, Nad99, PTB99, PJ97, Smo97, VBCBS05, Wal04a, dB92, vB05, BR96b, BC98, BE01, BD03, BT06, BR07c, CBPS10, DF02, MHS10, PAC08, BWT95, Dam99a, EJS98, Ste05]. Practices [GS00b, OSC^{+}08, STP^{+}10, SMBQ07, CT99, TDS94]. Pradesh [PSDS11a, PSDS11b]. pragmatic [Had96]. pragmatics [Gog94]. Prague [AM91c, BM05b, BTQ03, BBK12, BKL^{+}12a, BKL^{+}12b, BJ12b, CUDL02, CKMS05, De 07a, EV09, FKK04, FPCK11, FN12b, Gie10, GSS09, GV95, GRP10, GDGB12, HKR07, HS94, HK92, HKP10, HS95c, HMP99, HSH08, HdL12, JJP^{+}02, KMW03, KS01, KPP08, KLFK10, KNK08a, KNK08b, LD97, LMŠT01, Mac06, MP03a, MJ04, MSZ90, MST92, MRŠ03, MMP03, MMV04, Nes99, OPT01, PM04a, PM04b, PM04c, PM04d, PTRH05, PP12a, QBY^{+}12, RRBE09, SLH04, SKK04, Ste99, ŠPTM00, Tra04a, WH95, ŽL08, ŻR99, vWN99, vW97]. Praha [CK84]. Praque [HŽ07]. Prato [BH08]. pre [BG91b]. pre-conference [BG91b]. predicate [Dau03]. Prediction [CSM10, SYNC11, YGE12, Dob81]. Predictive [QCDM06]. PRELUDE [LT11a]. PReMI [CMM^{+}09a, GDP07, KMKP11, PBB05]. premières [GB77]. PREMO [DHM99]. presentation [WSL93]. presentations [VT97]. Preservation [IFL^{+}12]. Preto [SCR06, TDWG11]. Pretoria [WW03, YJKSW10]. Prevention [DPNM12]. PRIB [CNA08, DTMH10, KSG^{+}09, LWRdR11, RWA06, RSV07, SKSA12]. PRICAI [AIL12, FG96, HZ08, IS02, Kow01, MS00b, YW06, ZGY04, ZO10]. PRICAI’96 [AGT98, CRW97]. PRICAI’98 [LM98b]. Price [SRT06]. PRIMA [BK05a, BHH08, CS12c, DLW12, GGS09, Ish99b, KjHGG11, LS09, SS06b, YYI^{+}09, YY01, LB03, RWSS12, ZS00]. PRIMA’99 [NcC99]. Primality [Die04]. PRIME [CLS11]. primer [Bli87]. PRIMES [Die04]. primitives [BP95]. Princeton [AJRS03, BF12, CJRR07, GJRR11, GM08b, LHK02, TT95]. Principle [KjHGG11, Bur91]. Principles [BEGP04, DKW99, FSS06a, JTB^{+}05, KM04, KKdM^{+}07a, LGTB03, PMÓ06, Pau01, RWSS12, SKS^{+}09, BB96, GN90, ARS09, ABMS06, APW06, ALR11, BBT08, Ben06, Bes07, Bor94, BHM03, Coh10, Dec00, DG12, DLW12, EMT02b, FS05, Fre96, Gen09, Hig05, Jaf99, KZ97, Lee11, LMM10, MP98, Mil12, MR95, MNW98, Nad99, NW97, OS04, PGM98, PH04, Ros03, SSN08, Shv06, Smo97, Stu08, TTF07, Van02, Wal04a, Wal01, YYI^{+}09, ZKZ00, ZQ98, ŻR99, dRS01, vB05]. priority [EDD04]. PRISM [BB96]. PRISMA [Ame91b]. Privacy [AZ11, APP12, BS02a, BSN06, BN09b, BGW06, CLS11, CBCGA12, Dat09, DFS08, DFM11, DFT12, FHFL06, FHLP09, FHKQ12, FJ10, FKL08, FLP11, GKL09, GANACBR10, GANACL11, GANACBdCdV12, JMK09, JP12b, KLP04, KLS10, LPT07, Li11b, LPK05, MSS08, PH11a, PGD07, SNS03, STP^{+}10, SMCM07, SH10, SMS12, WPV04, Yal10, BCD00, BG05, DB98a, Fed01, FH01, MTW05, PS96b, PSNS99, San02, VPM97, VM01, AH10, BJSW09, BFMS08, BFJM09, BG07a, BG08, DM06, DG06a, DGDM^{+}11, DS03, Din03, DFT04, DFF06, DFS08, DFM11, DFT12, FHH11, FHW12, GA09, MS05]. privacy-enhancing [FH01]. private [Aso04]. privately [Aso04]. Pro [CGMP11]. Probabilistic [CPZ07, HS02, DG01, Kat99, DFKM08, Neu04]. probability [Rao95]. Problem [CCD^{+}12a, CCD^{+}12b, Han81, Han02, Hof03a, MA03c, MÁ07b, RJL^{+}08, RBB^{+}06, SCKR10a, SCKR10b, ZZ04b, Cas81, DSM94, EV96, EBSS98, Gra91, Kåg98, LPA00, MAB^{+}02, Oss99, SDR^{+}00, SM91, YBL^{+}04, Fen00, Fer93, God96]. Problem-Solving [MÁ07b, Oss99, Fen00]. Problems [AB11, AJRS03, BS06a, BJP12, BP09b, CKD11, CK12a, CJRT05, DP11, DJRZ06, FLP12a, GJRT01, HJRS99, JKRR04, KP10c, LMT10, PT08b, RR04a, SS76, VW07, vHH09, BJK06, BM05b, Bil97, BHR84, Chi79, FR95b, Fer96, FP96, Fer98, FP90, KK87, MM82, Thi00b, De 87, Koj91, Sud95]. PROBMIV [DG01, HS02]. Procedures [Ria08, Ria09]. Proceeding [BCT02a]. Proceedings [AvLR91a, AvLR91b, ADG^{+}08b, AHS11a, AHS11b, AS02a, ANS04a, ANS04b, ANS04c, AMSM^{+}09b, AKM10a, AKM10b, AvAS06, AvASD06a, AvASD06b, AvASD06c, APPE09a, APPE09b, ANS^{+}09a, ANS^{+}09b, AH03, AGS^{+}11b, AGS^{+}11a, AKL^{+}99, AHJ07a, AHJ07b, AHJ07c, AAH^{+}10b, AAH^{+}10a, AWA^{+}11a, AWA^{+}11b, AOM07a, AOM07b, ADGM12a, ADGM12b, ADGM12c, Ayk07a, Ayk07b, BBE01, BG91a, BBGS10a, BBGS10b, BBGS10c, BPAB07a, BPAB07b, BHH04b, BHH04c, BS11d, BS11e, BTQ02, BTQ03, Bau99, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BDIR07a, BDIR07b, Ber98, Ber02b, BBKW03, Ber00, BKL^{+}12a, BKL^{+}12b, BC85, BTBP^{+}10a, BTBP^{+}10b, BLKW01, BTCW10a, BTCW10b, BHR84, BFMR96a, BFMR96b, BFH98, BDGK00, BJKK02, BHQ10a, BHQ10b, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BvADS08c, BvADS08a, BvADS08b, BPSW06a]. Proceedings [BPSW06b, BLPW02, BGMST09a, BGMST09b, BCC99, BC96b, BC96c, CSPC09, CRJ11a, CRJ11b, CSCO00, CK04a, CK04b, CK06b, CK06a, CK10a, CK10b, CK12b, CK12c, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CPFD^{+}03a, CPFD^{+}03b, CFZ10, CZF10, CT03, Car03, CBNR02, CCD^{+}06a, CCD^{+}06b, CCP10a, CCP10b, CP98b, CP98c, CCD^{+}12a, CCD^{+}12b, Coe02, CRS10, CKW11a, CKW11b, CSA^{+}12a, CSA^{+}12b, CMPW12a, CMPW12b, DGT10a, DGT10b, DGSM11a, DGSM11b, DGM08a, DGM08b, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, DG02, DPB^{+}04a, DPB^{+}04b, DMLW11a, DMLW11b, DMLW11c, DCC^{+}10, DBD^{+}10a, DCC^{+}11, DBD^{+}11, DDI10a, DDI10b, DDI10c, DH04, DLŠ11a, DLŠ11b, DK02a, DK02b, DPMP^{+}11a, DPMP^{+}11b, DKP00, Ehr85a, Ehr85b, EV01, Ekl94a, Ekl94b, EP03a, EP03b, EW03, EL02a, EN74, Esp01, FC04]. Proceedings [FSdlPT11a, FSdlPT11b, FR01b, FMP11a, FMP11b, FMP11c, FLP^{+}12b, F^{+}12a, FLP^{+}12c, F^{+}12b, F^{+}12c, F^{+}12d, FLP^{+}12d, FDC12a, FDC12b, FMdGC99, FTZ08a, FTZ08b, FTZ08c, FTZ08d, FMC12a, FMC12b, FMC12c, GHJ06a, GHJ06b, GHJ06c, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, GF03, Gel11b, Gel11a, Gel12a, Gel12b, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, GG07a, GG07b, GG07c, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, GK01, GPB02, GLC^{+}11a, GLC^{+}11b, GPV00, GCC^{+}02b, GP02, GGK^{+}09a, GGK^{+}09b, Gue99, GHMV11a, GHMV11b, GHMV11c, GKN03, HLSZ11a, HLSZ11b, HMV01, HPV00, HSP^{+}12, HG98, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, HJW00, Ho12a, Ho12b, HDGK11a, HDGK11b, HYPY10a, HLI06a, HLI06b, HLI06c, HCS11a, HCS11b, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, IS02, IS12a]. Proceedings [IS12b, Jac07c, Jac07d, Jac07b, Jac07a, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, JK00, JKL02, JNR11a, JNR11b, JNHJ10a, JNHJ10b, JNH11a, JNH11b, JLS11a, JLS11b, JHC02, JNPV10a, JNPV10b, JNPV10c, JWbG^{+}06, JWG^{+}06, Jul98, KC11a, KC11b, KP00, KvETvdH10b, KP04, KWCW06a, KWCW06b, KWCW06c, KILW10a, KILW10b, Knu01, KvG98, KSDO06b, KSDO06a, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KM00, KP15, KNK08a, KNK08b, KKGT03c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, LSS02, LSW00, LN01a, LNS06a, LNS06b, LM98b, LTL^{+}11a, LTL^{+}11b, gLPZ^{+}12a, gLPZ^{+}12b, LBP06a, LBP06b, LBP06c, LBP06d, LLC09a, LLC09b, LFJI10a, LFJI10b, LJS^{+}10, LSTZ12a, LSTZ12b, LN03b, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LDXZ10a, LDXZ10b, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, LHJ08a, LHJ08b, LHJ08c, LZK11a, LZK11b, LZK11c, Mag02, MF11a, MF11b, Mar11a, Mar11b, MSS12, MS10b]. Proceedings [MS10c, MS12b, MBMS07a, MBMS07b, MN02b, MZ98, MS02a, MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MDH09a, MDH09b, MDH10a, MDH10b, MDH^{+}11b, MDH^{+}11c, MMO^{+}11a, MMO^{+}11b, MT99a, MAFS08a, MAFS08b, Mey01, MKZK10b, MKPZ12a, MKPZ12b, MdPA98, MSA99a, MSA99b, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, MKR^{+}04, MS00b, MF02, MAFSS04, MMK98, MSY00, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, NNS06a, NNS06b, NHJ04a, NHJ04b, NHJ04c, NLŚ10a, NLŚ10b, NKJ11a, NKJ11b, NHJ12a, NHJ12b, Ode04, ORS89b, OZ01, ORB^{+}09, OF15, PR01, PM04a, PM04b, PM04c, PM04d, Pal00a, PCN10a, PCN10b, PCN10c, PCN12a, PCN12b, PCN12c, PSDS11a, PSDS11b, Pat00, PTB99, PRH10a, PRH10b, Pol99a, PT91a, PT91b, PR98b]. Proceedings [QGZ03, QLK^{+}10, QLK^{+}11, RBGS04, RDPMA^{+}11a, RDPMA^{+}11b, RR03, RT00, RCS10, Rov98, RKS^{+}12a, RKS^{+}12b, Sak01, SS09a, SS11c, Sd98, SCKR10a, SCKR10b, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, SBC^{+}94, SvADS07a, SvADS07b, SvADS07c, SvADS07d, Shu11a, Shu11b, SAG10a, SAG10b, SYP^{+}05, STDH02c, STDH02a, STDH02b, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, SS07c, SS07b, SS09c, SS11d, SPB00, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, SCB01, SRL12a, SRL12b, SRL12c, SZT^{+}08a, SZT^{+}08b, TST10a, TST10b, TSCW11a, TSCW11b, TCHB12b, TCHB12a, TSJ12a, TSJ12b, TKCW11a, TKCW11b, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, TNVDF06, VPBŠ03, VRHJ09a, VRHJ09b, Ver00a, Ver00b, VDÉ^{+}12b, VDÉ^{+}12a, Vos03, WLY05a, WLY05b, WLY05c, WYZ06a, WYZ^{+}06b, WDGL10a, WDGL10b, WYP12a, WYP12b, WWD99a, WWD99b, WMB10b, WMB10a, WD10a, WD10b]. Proceedings [XCHZ11b, XCHZ11a, XSA^{+}12a, XSA^{+}12b, XHXL08, XLHX08, YKKZ07a, YKKZ07b, YHR^{+}09a, YHR^{+}09b, YWG04a, YWG04b, YHZ09a, YHZ09b, YHZ09c, YKU11a, YKU11b, ZYRP10a, ZYRP10b, ZRP^{+}11a, ZRP^{+}11b, ZGY04, ZLK10a, ZLK10b, dBNT87a, dBNT87b, dBM10a, dBM10b, dSADM07a, dSADM07b, dPMA98, AK90, AM91b, AM91a, ADG^{+}08a, AM97, AMSM^{+}09a, ADJ^{+}01a, ADJ^{+}01b, AF02, Ame91a, Arn90, Aus86, BK90, BM99a, BYGP95, BPP91, BB90, BSW95b, BE87, Bec79, Bes93, Bet83, BCI85, BHCL87, BJ86a, BD97b, Bis88, BH77, BZ83, Boh75, BBR88, BBR89, BBR90, BM98, BS03b, BH95, BHT97, BVNW87, Bra80, BRR87a, BRR87b, BHN97, BN98a, BN98b, Bur90, Cea76a, Cea76b, Chi79, CL90, CK84, CJK88, Cle92, Coi96, CDW86, CW88, DN88, Dav89, Dem80, DT89, Des86, DO89a, DO89b, Ehr87, EEK^{+}87, EKLM87, EMW94, EK81a, ES92]. proceedings [FK87, yF75, FJ92, FJL97, FG96, For88, FZ86, Fri93, FTF88, FTF89, FTF91, GJ86, Gan88, GJ93, GW96, Gec81, GB97, GM89, GK90, GN88, Gol90, GH96, Got83, GAY86, Gru77, GC81, GRW86, GPV88, Hal94, Han81, HHJ^{+}86, HAM95, HS76, HL91, ILR95, Jon80, Jon90, JP94, Jos88, Kah79, KG74, KL88, Kar77, Kar83, KHJ05d, Kir96, KB92, LS95, Len97, LB78, Lor01a, Lor01b, LD05b, LD05a, Mad92, MB82, Mak86, Man75, Mar75, MdlBP05a, MdlBP05b, MLM90, MM89, MP95a, Maz76, MW97, MC89, MS96c, MNS95, NDB94, NDF94, NS87, Nic75, Nic86, Nie96, Nie93, Nie94, Nol81, ORS89a, Odl87, Ohl93, Olt95, Pap95, PR84, Pic86, PJ97, Pom88, PR96, Rao92, Rei88, RM97b, Rob74, Rob80, RW86, RBW95]. proceedings [SS83, Sal83, Sam76, Sch93a, Sel86, Sim92, Sta82, SW75, SvASD05a, SvASD05b, SvASD05c, TPB79, Tob80, VY93, Wad86, Wad87, WF77, Wil86, Win78, WL93, YK82, vF91, vdV84, ÅBE12, AH00, AO02, AAW04, AK10, AG01, ACvLdR03, ACdR^{+}08, AD05, APFV09, ARS09, AK07, AP09, AL11a, AGB^{+}11, AKZ01, Abe06, Abe10, ABB12, APR^{+}06, ACN^{+}07, AV99, AM10, AJNJ12, ABS01, AMS06, ABMM07, Abr07, ALS^{+}11, Abr01, ABD^{+}09, AI06, AB11, ASS10, AJRB09, ARSB09, AST05, AFvZ02, AR99, ABB^{+}11, AAA^{+}11, AT02a, ABK^{+}09, AFP^{+}10, ADDL08, ACL12, AW12, AYV^{+}11, Agu12, ABG09a, AW01, AB12, APS08, AVP07, ACD^{+}07, Akm01, ABCWM99, ASE^{+}07, ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12, AG98, Alb99, AR04, AS03, ADP09, ABBC10, AD98, AL04, AGL^{+}07]. Proceedings [AE05, AD06, ABB^{+}10a, AF11, AAB^{+}06, ARC^{+}07, AH02, ADR07, ABB10b, APSZ02, AV08, ACJ09, ABB99, AFM01, ABMH08, AJB09, AS11b, ANF08, ABR09, AR10, AL03a, AP04b, AP04a, APBP07, AMGJ12, AL03b, Ama08a, ACR07, AC11, ACK00, ASLM09, Amb01, AYBH^{+}06, ABD^{+}99, Ami98, AL01, ATM09, AP10, ASR^{+}07, AMRŚ08, ALPH10, AFG^{+}11, AFS02, AHK07, AFGvE10, AFGvE11, ACH00, ABM02, AFL03, AH98, ABH^{+}10, ADMH04, ADM^{+}06, ADM^{+}08, ACL98, AMCL02, AYB^{+}09, ABK12, ADL05, ATO^{+}09, APHA11, ALR11, AIL12, AB04, AP11, AT12, AJ11, AT02b, AGI^{+}06, AMPT09, AT02c, AS07a, AZ11, ABM05, AF06, AHW07, AM98, AF99b, AM00, AJS00, Arm02, ABD08, AI98, AJRS03, ATC^{+}09, AN07, AN09, AKG06, AR98, AP07, AHMT08, Asa99, Asa06, AiNOW11, AB06, AB03, APP12, ABT04]. Proceedings [Asp98, ASAM07, ABRZ06, AH07, AH10, ALZ11, Atl08, AD11, AJ01b, ACL^{+}04, AL06, ACJ08, ACR12, Auc98, APM08, AM05, ACR^{+}08, ACD^{+}10, AKTT12, Avg12, AKSO06, ADS09, AGCP^{+}11, ADK04, Ayk09, ALL^{+}10, AE06, AS07c, AKR^{+}09, Baa03, Baa07, BV02, BM03a, BVO10, BG10a, BW05a, BE11, Bac00, BN09a, BO00a, BC09a, BBCL08, BLPW03, BBdB12, BYCC03, BYdVZ^{+}12, BS11a, BA06, BH06a, BMS^{+}08a, BCBF08, BBT08, BHL^{+}03, BCB99, BMK08, BMK09, BDK10a, BKH11, BLS06, BST01, BST03, BR03, BJ06, BSC12, BB07, BM05a, BMUV10, BKS99, BS11c, BR12a, Ban98, BCD^{+}03, BDZ04, BLO^{+}07, BLW09, BW11, BSZ12, BGG^{+}12, BA99, BP04a, BBS07, BKA11, Bar08a, BBB^{+}07a, BP12a, BH06b, BFH07, BCS09, BCC06, BPCS06, BA07, BMPM09, BBH^{+}10, BBA12, BFF^{+}10, B^{+}12a, BGK12, BM05b, BdB08]. Proceedings [BH10, Bar11, BPS11, BLS12, BG09, BBAB^{+}07, BL06a, BP07, BW03b, BR04, BGP08, BGP11, Bat01, BCT02b, BS02a, BSN06, BCL03, Bat09, BW10a, BGV01, BMP00, BMP01, BBP04, BPW06, BCE09, BCM08, BD12, BHHK08, BS06a, BPV04, BPR08, BWAM09, BH08, BBLT06, BDL08, BBK12, BB99, BL05b, BCRB^{+}11, BdQ11, BJP12, BH06c, BPR02, Bel03, BMR^{+}04, BL08, BHRU09, Bel00, BP11, BCC^{+}10, BAHH07, BE08, BC10a, BGKY08, BEP04, BBB07b, BDCM04, BS03a, BCG^{+}07, BCKR10, BCST99, BDG04, BKR09, BB01a, BG11a, Ben06, BP03, BLJ08, BDW08, BMSHW09, BFBS11, BF09, Ber10, BMS08b, BJM09, BNUW09, Ber08, BLKP08, BF12, BFGL10, BKH^{+}09, BL10a, BC11a, BGPTF10, BF11, BC06, BDSW09, BLB^{+}03, BGF06, BSTL07, BCP^{+}04, BRL08]. Proceedings [BDH^{+}99, BCP99, Bes07, BG11b, BND01, Bha00, Bha02, Bha05, Bha07, BKW08, BKW09a, BW10b, BB03, BM10, BFG^{+}12, BN04, BG06, BS01a, BG03b, Bih03, BG12b, Bil98, BdBK03, BD98, Bir12, Bis02, BV11, BL07a, BR02, BF00, BBRM98, BBZ00, BSS11, BV12, BSBM05, Bla03a, BvDMST11, BMS04b, BTPPS06, BTPPS07, BTBP^{+}08, BTPPS09, BTKP^{+}11, BTPP^{+}12, BP01, BMSW12, BBRS05, BBC^{+}08, BBD^{+}09, BBR^{+}10, BS02b, BC10b, BKV10, BL03, BMO02, BPS06, BHHJ11, BR12b, Bod05, BL06b, Boe01, BJS99, BGPW05, BKLP11, BKLR12, BU01, BM02a, BDS04, BMW10, BTCW12, BG10b, BDK10b, BTZ^{+}10, BO00b, BKS01, BI04, BKS05, BJ12a, BTF02, BBR07, Bon03, BGP00, BDDM12, BJ07, BB00a, BNS00, BGR03, BBBB08, BG08, BGPR10, BH02, BF06, Bos01, BK04a, BL10b]. Proceedings [BM02b, BHJ^{+}11, Bos00a, BL02a, BE07, BF10, BM09a, BC10c, Bou11, BMCCM11, BBMSD09, BKM08a, BHL11, BCK10, BEGP04, BEG04, BJ01b, BIMW12, BSB^{+}99, BZ08a, BP09b, BEM06, BAP^{+}05, BCD00, Boy01, BM03c, BN09b, BC11b, BKLS11, BH05, BR09, BS01b, BL07b, Bra02, Bra09a, Bra99, BRT11, BTT12, BW00, BL01, BLV02, BAH09, BvdHPP09, BCKT05, BCF06, BH00, BKZ05, BNZ06, BZ09, BHV11, BJ01c, BW01, BC09b, BHH^{+}06, BFvdM10, BKW06, BKNS06, BMS01, BD04, Bri93, BBH08, BUHS08, BGR08, BL02b, BRP09, BFMS01, BP12b, BR98, BJ12b, BB02, BV05b, BD11, BK04b, BM11b, BSS00, BCR03, BSL^{+}09, BHM03, BG07b, B^{+}12b, BAWH00, BMC08, Buc85, BM06, BCF^{+}02, BD08, Bue04, Buh98, BHH08, BOH^{+}12, BH11, BZ08b, BLVV07, BL06d, BDFS04]. Proceedings [BkHJ^{+}04, BLS02, BRS09, BPS02, BHLP07, BS11f, BKO03, BKO04, BFKO06, BFK^{+}07, BFK^{+}08, BFC^{+}09, BL11, BB00b, CPS05, CV11, CC04, CAD^{+}02, Cag00, Cag02, CCL06, CLKL09, CHKL10, CV07, COBV11, CFI06, CD10, CG07, CBGCCZ12, CYM^{+}11, CP98a, CBC^{+}02, CIW06, CDP02, CDV03, CCD04, CDP^{+}06, CCF^{+}08, CCD^{+}09, CHM^{+}10, CKPR11, CLM03, CL08a, CKS04, CPR^{+}10, CK08, Can08, CV04b, CYGL04, CSJD06, CLWC08, CE01, CHD07, CDCW09, CZ02, CS07, CP10a, COT98, CNP^{+}08, COMTLB12, CMT02, CBF09, CP10b, CT06, CFM09, CF00a, CGS01, CVTP12, CJ08a, Cas09a, CVLL12, CL01, CFGN11, CIT11, CFH^{+}12, CD09, CDGW10, CD07, Cav85, CL07, Cer05, ČGH^{+}11, CP11, CC07, CD00a, CGP02, CCPP12, Cer07, CVV08, CCJ07, CCK^{+}08, CT07, CSM10, CM12, CG02, CGBZ08, Cha02a, CCL02]. Proceedings [CL06a, CWC^{+}07, CKK^{+}09, CHCM11, CZG^{+}08, CGKM11, CMMR12, CJRR07, Cha02b, CPC^{+}99, CH10, CWZ^{+}11a, CWYC12, CDPT06, CMM^{+}09a, CSR08, CL10, CW09, CLL^{+}09, CGL^{+}11, Che99a, CLY02, CCH02, Che03a, Che03b, CCM^{+}04, CMZL04, CL06b, CWY^{+}06, CDLZ06, CJ08b, CRW08, CMS08a, CYCL09, CC09a, Che10, CChL^{+}10, CTYG10, CTS10, CWZ11b, Che11, CEK11, CFLR12, CC12, CP06b, CLZ06, CKLN10, CEQ^{+}11, CNA08, CWL01, CCdS03, CFK^{+}12, CvSW04, CM03b, CHCA09, CY12, Chi04, Chi03, CJ06, CH09, CPC07, Cho12, CQSS09, CW00, CRD08, CKH10, CCAS12, CCBW11, CDCJL08, CH99, CS05b, CHAL04, CDY^{+}11, Chr99, Chr00a, Chr00b, CKM^{+}09, CHA03b, CM06, CI98, CM04a, CW99, CW06, CM09b, CS12a, CP10c, CP99, CLB11, CPPB06, CCM08, CA10, CG09, CU08, Cle99, CGS07, CS00b, CFG^{+}11, CH98, Cod01]. Proceedings [Coe98, CGK^{+}12, CAZ05, CSTD08, CBT07, CS02a, CAB10, Coh10, CMS11, CÓ11, Coi99, CTE^{+}06, CK01, CGS02, CSAH09, CBF01, CMP07, CDSS07, Con00, CS01a, CNmWH^{+}05, CGGO03, CS02b, CP07, CLS07b, CK07, CKOB11, CDL12, CYBF06, CAP08, CWO^{+}09, CY09, CAdB01, CBW99, CEKR02, CEM^{+}06, CKC11, CHK06, CTS07, CR04b, CF99, CCSW12, dRCVT11, CDFP11, CM04b, CD01, CDND11, CvH07, CC09b, CSRS08, Cou01, Cou03, Cou05, CM11, CMB^{+}08, Cov99, CM10, CPR12, CS01b, Cra08, Cra12, CL08b, CP02, CBBS09, CPP07, CMM01, CDS09, CGvR02, CR05b, CFS06, CSSW04, CGB11, CFL10, CFRR10, CRS00, CH01, Cro03, CDT11, CDA^{+}06, CVÉ07, CMS08b, COB06, Cue03, CHR10, CBCGA12, Cup00, Cur09, CFB11, CF00b, CD11, CD12, CT99, COB^{+}08, D’H10b, DW03a]. Proceedings [DN07, DSZ04, Dai07, DL06a, DO02, DH07a, Dam99b, DL06b, DF01a, DVVA08, DB02, DD03, DSDS04, DG04, DPLW08, DU09, DH01, DG06b, Dat09, DSZ^{+}11, DFUR11, DMO98, DF01b, DFR02, DM01a, DSB04, DMS04, DJAP07, DG07a, DCFR12, DV10, DdQ10, DB98a, DW07a, DEKR09, DG01, DAB01, DKS^{+}07, DKC10, DSB10, DLNU11, DC12, DMR01, DR11, DFM04, De 07b, DPZ12, DY06a, DMB^{+}11, DKK08, DB04, DGJ^{+}05, DHRT10, DZ10a, DBC^{+}10, DMW00, DRDKE11, Dec00, DIMV09, DFMV10, DIMV11, DMV12, DPV11, Deg03, DG09a, DG12, Deh99, DST01, DSS03, DSZ07, DGST09, DIS11, DKM^{+}05, Del00, DF11, DS01a, DP11, Dem07, DH11, DB98b, DQBZ02, DBPZ05, DG07b, DWWL09, DHX09, DBB^{+}08a, DsSL^{+}07, DNS09, DV07a, DFG^{+}12, DGLT12, DV07b, DS98, DPW04, DH10a, Des02, DBM09, DSM03, DKLT05, DZ03]. Proceedings [DSV01, DAC^{+}12, DPP09, DPT12, DY10, DB09, DOS99, DKLS04, DD10, DVM^{+}11, DVV07, DN09, DC00, DJJ12, DTMH10, DR02b, DS08, DBHS10, DZGS06, DSZ10, DLdS04, DMB^{+}07, DGN01, DdCF98, DLGJ10, DW10, DPT08, DSL08, Dol06, DHO08, DO09, DCFY10, DIP12, DFT04, DFF06, DFPS06, DFS08, DFM11, DFT12, DPSU11, DP12, DT06a, DGG^{+}10, DTW03, DW03b, DLW^{+}07, DDI09, DZ10b, DLM99, DLO03, DPNM12, DBH01, DDS02, DGB^{+}06, DBB^{+}08b, DBD^{+}10b, DBB^{+}12b, Dor02, DGJ09, DF06, Dov05, DFD04, DXZ07, DTF98, Dro08, Dro09, Du00, DHP09, DFW^{+}11, DSS^{+}07, DP06, Dub09, Due04, Duf07, Duf09, Duf11, DHH^{+}08, DRS08, DH08, DA01, DBB^{+}08c, Dun12, DH07b, DLM07, DLJ12, DT06b, DFGN12, DFS06, DKW07, Dwo06, Dwy01, DL07, Dyc00, DF99, ET01a, ENP02]. Proceedings [EOE^{+}07, EKS10, EB04, ERW99, EM03, EHKP05, ED06a, EBT11, EK09, EM02, EGMW11, EM08, EGO09, EF02, Ehr98, EFG^{+}09, EHRT08, ERRS10, EEKR12, EBSS98, EdS08, ES10, EFT01, ES02, EL04, EK12, Ekl04, EH08, EKR10, ECB06, EHH12, ERK07, EM06, EK10, ELNM08, ELN^{+}10, EML^{+}12, EMT02a, EMT02b, EHM11, EOR06, Emb98, ES00a, EN06, EW04, ES03, EOSW07, EYX^{+}05, EBT07, EKO^{+}04, EF12, ELS09, EWZ11, Ern07, EV07, ETG02, EF03, Ési06, EAES^{+}10, EL02b, EM10, ERMS06, Est98, Est99, ET06, ES00b, EKM06, ED06b, EKS00, ER12, FRvKS10, FHJ^{+}02, FT11, FSA^{+}07, FL00, FWY05, FPP11, FC98, FGH98, FK10, FD03, FMC04, FS07b, Fed01, Feh11, FR98, FR00, FR11, FKP99, FLY06, FWZH06, FS99a, FKB02, FV10, FB12, FL08]. Proceedings [Fer08, FR09, Fer98, FMS10, FLMG10, FIAP01, FFL^{+}12, FM04, Fes10, FI08, FKK04, Fic03, FMW01, FK02, Fie02, FV09, FOGE^{+}09, FZJ01, Fin99, FLP12a, FHFL06, FHLP09, FHH11, FHW12, FHKQ12, FS99b, FTAZ06, FDL10, FW07, Fis09, FBM12, FvdHKL06, Fis08, FvdTDG12, FSN06, FHT05, FHY08, FBV11, FK12a, FB09, FT04a, FX08, Fle02, FG06, FLF^{+}08, FMN99, FKS10, FR99a, FSV09, FK12b, FCLL02, FKW08, Fon03, FRL07, Foo99, FP08, FJ10, FYM12, FGK^{+}11, Fos99, FHP03, FILP06, Fos02, Fra00, FR99b, FKBP11, FKBA12, FW09, FKM07, FRSH05, FMNS06, Fra01, Fra99, Fra04, FHW08, FGK^{+}10, FG10, FdS12, FSTS09, FCD^{+}04, FT04b, FLL03, FCLG04, Fre01, FM99, FKS07, FNGW08, FGTZ08, FLR11, Fri01, FMRW09, FHP11, FSP09, Fro02, FD11]. Proceedings [Fuc98, FK09, FS06, FN12b, FKL08, FLP11, FSS06a, FSS06b, FKLY08, FTC^{+}10, Gab13, Gac02, GGD09, GP07, GP09a, GP11, GPAS^{+}06, Gal07, GP08a, GN12a, GTT04, GCB^{+}05, GBH11, GLP12, GE08, GM00, GG05b, GMV06, GMV07, Gan99, GVM99, GWY99, GJ09, GLSY10, GLW^{+}12, GH03, GLMP07, GMO09, GD10, GMS03, GY04, GG09, GWK09, GB99, GRT02, GH08, GVT08, GL10a, GNR07, GFV00, GJT03, GLZZ09, GGK06a, GLCT10, GKC^{+}12, GKB^{+}08, GPAD08, GT03, Gel01, Gel02, Gel03, Gel04, Gel05, GdATM05, Gel06, GRG06, Gel07, GM07a, Gel08, GM08a, Gel09, Gel10, GLP99, Gel99, Gen09, GRB^{+}07, GM03, GFV09, GMV09a, GKMV10, GKMV11, GKMV12, GK06, GL10b, GN12b, GS09a, GM04a, GDP07, Gia07, GBC^{+}08, GBC^{+}09, GVB12, GS00a, GH07a, GM11a, GO11, GM12]. Proceedings [GAA06, GO09, GN12c, GW09, Gie10, GR12, Gie05, GH10, GMBM05, Gil10, GS09b, GHI^{+}12, Gir99, Giu98, GZR02, GK12, GH09, GSH^{+}04, GMI06, GMUW12, GM06a, GP09b, GP98, GJRR08, GJN10, GJRT01, GRK^{+}10, Goe98, GW11a, GCSR07, GZ09, GA09, GJRR11, GKW02, GMS06, GF12a, GPE05, GPYD09, GHSY06, GG10, GdlP99, GTVC06, GCM08, GP12, GW98, GQ11, Gor03, Gor10, GLN01, GPS03, GLS05, GZSC07, GW06a, Gór06, GHC^{+}06, GH12b, GR06a, GGS99, GM06b, GHP09, GS10, GK09, GBMS11, GS00b, GM04b, GZ06, GMS12, GSW06, GYA^{+}06, GJS08, GHH^{+}06a, GL08, GTY03, GSS00, GL03, GHH06b, GS08, GPT10, GM11b, GN08, GAM^{+}11, GSS11, GW12, GH04, Gru98, GH07b, GMV09b, GN98, GRP10, GDGB12, GV09, Gud08, GMV12, Gue01, Gue04, GP09c]. Proceedings [GG02b, GPP09, GW11b, GJTS06, GY03, GYD^{+}06, Gup99, GM08b, Gup10, GM10, GJRS12, GKOT00, GPL99, GW00, HNH^{+}09, HP02a, HOC07, HM11a, HRZK02, HP12, Had03, Hae99, HK04, HW06a, HRK^{+}07, HKR07, HP99a, HZ05, HR06, Hal09, HG02, Hal00, HPTV05, HCT02, Ham08, HT09, HMT10a, HT11, HHMT12, Ham00, HS07, HSJ07, HTW02, HY12, HRdV11, HP99b, HV03, HMS09, HWW^{+}10, HKB99, HAB02a, Han98, HMT10b, Han07, HM12a, HS06, HKP08, HKL07, HHLS08, HRC06, Har07, Har09, Har11a, HS08a, HMTH10, HK98a, HG00, HK05a, HKI08, HZK08, HH10, HZ10, HS08b, HMP08, HBS00, HLMW06, HM12b, HKS06, HS00, HHP09, HLKX12, Hed03, HNM02, HKP10, HU08, HLW04, HMW^{+}08, HJ12, HW01, HJMP06, Hen08, HA02, HWG10, HK10a, Hen06, HS98a, HK01]. Proceedings [HRB12, Her00, HV06, HS02, HP06, HP02b, HC05, HM02, HC11, HMM12, HIS05, HAB98, HHJZ12, HBE07, HHW01, HT06a, HPP06, Heu98, Heu00, HP09, HN12, HK11, HA00, HSS07, HW09, HRSS08, HKLP12, Hir98, HL10, HMP99, HK10b, HCL05, HZ08, HSKB09, HJRS99, HAB02b, HW07, HdQ08, Hof03b, Hof11, HW03, HLW08, HKT12, HSB^{+}10, HB11, HKP11, Hol07, Hol08, HM09, HS11, Hon01, HS98b, HNF08, HTMC09, HM01, HRM^{+}10, HVT12, HRRY08, Hor00, HMG03, HSL^{+}99, HSH08, HCDL09, HH02, Hor02, HY03, HT06b, HNS^{+}12, HMPZ09, HS98c, HKV10, HO98, HK98b, HYMZ11, HV98, HRA02, HW08a, Hu09, HLCZ11, HSSV12, HM12c, HdL12, HC09, HH03, HZH05, HSL06, HHL07b, HHL07a, HKM^{+}07, HWLJ08b, HWLJ08a, HXC^{+}08, HYP^{+}09, HJL^{+}09a, HJL^{+}09b, HZBF10, HZGZ10, HJBF12, HMJG12]. Proceedings [HW08b, HL00, cHPkC^{+}07, HS99, HMT10c, HKH10, HLFS12, HS08c, HJ00, HS03, HP99c, HMCL06, HM98, Hus01, HS98d, HK07, HSTU99, HU05, HST07, HSVZ10, IdKM06, IKO03, IZ02, ID03, ID06, IY06, IR08, IKR00, ICSY07, IT06, IAC06, IZ98, IZ99, IZ00, Ima99, IR07, IMY^{+}07, IPRO08, IL01, IFGH10, IFL^{+}12, INR07, IS98, IAL^{+}07, IK12, Ish11, IS08, IT08, IZP09, IN11, JZR09, JNS05a, JP05a, Jaf99, Jäh99, JSV11, JOD98, JM05, JL08, JM11a, JYZ04, JLY03, JN10, JRL09, JM07, JM10, JR98, JLV02, JKRR04, JSWhT11, JS01, JT09, JO99, Jay99, Jay04, JFBH08, JM02, JZ03, JS08, JSW03, JL00, Jen01, JJP^{+}02, Jen03, JP04a, JPD04, Jeu98, JCC^{+}12, JP03, JDL11, JKN^{+}12, JSK10, Jha10, JM11b, JS11a, JWH06, JP09a, JFT11, JDAW12]. Proceedings [JPXS04a, JPXS04b, JGXC04, JRPP07, JM03, JV06, JK04, JLW07b, JMO09, Jon00, Jon01, JP05b, JP06, JP07b, JP08, JP09b, JP10, JP11b, JP12a, JVA04, JTB^{+}05, JMK09, JC12, Jos00, JMP12, JS11b, Jou09, Joy03, JQ04, JMO10, JMT11, JK06, JK05b, dCJB12, JN12, Kåg98, KSG^{+}09, KS09a, KG12, KSHS01, KPS12, KM11, KMTW03, KP06, KPK05, Kam99, KMT00, KWA01, KWA02, KMW03, KMW04, KC07, KC09, KKGK11, KS04a, KM08a, KJR05, KKKS11, KLZ07a, KLZ07b, KCYL08, KR10a, KPSW10, KL07a, KSS08, Kap10, KMMN11, KvETvdH10a, KR02a, KR04a, KJ09, KS12a, KWW04, KTH09, KS03b, KMD02, KV04, Kar09, KL07b, KMC^{+}01, Kat99, KS02c, KK11a, KK11b, KGL04, KLP04, KLB^{+}06, KLS10, KY07a, KWC^{+}12, KKMW07, KP10a, KLT^{+}12, Kay03, KLM^{+}07, Kei09, KOL^{+}04]. Proceedings [KTC^{+}05, KS01, KMF06, KGRD10, KS03c, Ker01, KBP07, KMS^{+}11, KLP10, Kia11, KSB10, KMSB11, Kil01, Kim01b, Kim01a, Kim02, Kim05a, KS06, KT06, KGV^{+}09, hKA10, hKhLKŚ10, hKAS^{+}11, KSW05, KjHGG11, KK98, KR00a, KR02b, KJB09, KR00b, KR01, KN03, KS02d, KSUZ11, Kla12, KY07b, KPS01, KŽ04, KGG^{+}11, KO11, KW98, KSW99, KK00, KZ01, KOS02, KOOL03, KOK04, KRP06, KHPS07, KPP08, Kno11, Knu99a, Knu02, KLLP12, KP01a, Kob06, KW07b, KP99, KP01b, KNP01, KS04b, KFW04, KKM^{+}11, KH02, KKdM^{+}07a, KKdM^{+}07b, KRRBV07, KK09a, KHL10, KK11e, KW07c, KHW03, KCMO11, KLPP03, KPK^{+}10, KC02, KKS10, KvdM00, KFMM06, KRH09, KV08, KK09b, KFLS07, KBH03, Kos98, KY11, KI12, KKMN07, KS10a, KS12c, KYT01, KHI06, KHS07, KMMT12, KGS08, KU12, KFM07, KHK^{+}08, KVMH09, KP09]. Proceedings [KR10b, Koz04, Kra03, KS04c, KU06, KN09b, KP05, KLN07, KO07a, KKL12, KKDV02, KT11, Kra99, KLFK10, KJ10, KSHM09, KR02c, KO07b, KP10b, KPR07, KP11, KOS07, KKH07, Kro00, KS12d, KLC07, KASS07, Kru08, KPRS07, KRSR11, KSTP99, KNP06, KK07, KP10c, KU01, Kuc11, KU09, KWT03, KF12, KSGT10, KR04b, KH12, KV11, KS08b, KH09, Kul09, KKGT03a, KKGT03b, KS10b, KR12, KMMP12, KJS01, KR06, KSG^{+}07, KMS10, Kur07, Kur09, Kur11, KLT09, KMR12, Kut98, KPW99, KCG09, KL02, KS07c, KMKP11, KŚHM11, KDWW10, KFfKL11, KS10c, LLT07, LH11, LBNF08, LLS00, LO02, LCD^{+}09, LGK^{+}04c, LC06, LZL11, Lai03, LMR^{+}06, LJR^{+}10, LOX99, LPT07, LK10, LM98a, LM06a, LLW06, LV07a, LLMT04, LSW98, LPOA^{+}06, LKD08, LLCL05]. Proceedings [LK99, Lau00, LGTB03, LGBT03, Le 02, LZ08, LOR10, LT00, LBP00, Lee01, LMW01, LV02, LB03, Lee04, LLWL04, LSgL^{+}06, LTW06, LL07, LKK^{+}07, LCHS08, hLhKcFŚ09, LLPH09, LSL^{+}09, LN09, LBM^{+}10, LYBU10, LCY10, LMMN10, LW11, LHŚ11, Lee11, LHKŚ12, Leh11, LWDM12, LTÅ^{+}11, LHH10, LM06b, LdQ07, LRG04, Leo04, Leo09, Leo00, LG99, LO98, LNE06, LVP04, LM09, LwCM00, LPC^{+}11, LW09, Lev98, LSY^{+}06, LPZ12, LSE02, LSHB07, LV06, LPH09, LLT^{+}04, LWF04, LZhL06, LYT06, LFIM07, LLIH07, LJJG07, LZS^{+}08, LSYO08, LKZ^{+}08, LHDD08, LFP^{+}09, LMA09, Li11a, Li11b, LCB12, LPR12, LNW^{+}12, LEP^{+}10, LMT00, LN03a, LS06a, LP10a, LFK^{+}02, Lin07, LTW11, LXH^{+}12, Lin12, LDTU04, LRL98, LJ00, LP10b, LPH00, LNS11, LM03, LYL06, LSK08, LKT08, LR05a, LMV98, LYhL^{+}01, Liu01, LH06, LMA08, LWYN09, LR09, LBDW09, LLWL09, LC11, LSIT11, Liu11, LMC12, LLTY12]. Proceedings [LS04b, Llo00, dVLKR^{+}08, LM04, LPR08, LMT10, LPA00, LPZ08, LN04b, LST^{+}12, LWRdR11, LLMR09, LO10, LPK05, LSF07, LT11b, LHK02, LNSS11, LGM11, LZ00, LZ05, LMM10, LRS98, LM98c, LR05b, LM08b, LS06c, LS12b, LTT07, LKK10, Luo04, Luo06, Luo07, Luo08, Luo09, LSW^{+}10, Luo10, Luo11, Luo12, LMB08, LK00, LS10b, LHH11, LIH99, LCDT08, LDC^{+}09, MJYT06, MZ07a, MRN07, MCA08, MOP11, MOH^{+}12, MOP^{+}08, MRC12, MPK04, MT09a, MDY^{+}10, MDHB11, MS12a, MCK^{+}06, MKPV05, MCKB06, MWYF10, MPV12, MTT98, MT09b, Mag03, Mag98, MSB12, MP98, Mah04, MPP08, Mai00, MBG12, MMV06, MT09c, MRvM07, MMM01, MP03a, Mal02, Mal08, MJ04, Mal99, Mal01, Mal03, Mal07, Mal09, Mal11, MZ99, MBT^{+}03, ML04, MSP^{+}09, MST12, MZ07b, MSZ08, MNZ09, Man09, MS10a, MN02a]. Proceedings [MPS06, MS03a, MSS11, MPdSM07, MM08, MT04a, MM01, MY01, MR01, MRŠ03, MMP03, MMV04, MSZ09, MVL11, MEAC04, MdRIR08, MBH^{+}09, Mar01, MSS07, MVCM10, MW11, MB01, Mar03, MNW01, MA03a, MOFM10, MSW07, MTOK06, MTCOK10, MTCOBH11, MBVTY11, MDY07, MRZ09, MWZ10, MMDN12, MMP07, MKT98, MT07a, MMOS99, Mat01, MM03, MM07, MM09b, Mat09, MSWZ06, MTO99, MF08, MW03, ML11, MZCC09, MP09, MP03b, May01, MSAdAS05, McA00, MBP^{+}11, MG07a, MT04b, McG08, MW09, MB06, MPvH04, MEB05, MNPW07, MO08, MTP^{+}02, MTS^{+}03, MT^{+}03, MTH08, MHD09, MDH10c, MDH11a, MZC06, Mei08, MT08c, MRSV07, MCA^{+}07, Mel07, MH04, MSS03a, MS02b, Men05, MSW02, MN98, MAB^{+}02, MHA09, MR07a, MM10, MH11, MR08, MF11c, MA11]. Proceedings [MM04, Mez11, Mic10, MK03a, MS99b, Mid01, MKZ02, MKZB04, MKZK06, MKZK08, MKZK10a, MR07b, Mil12, Mil00, Mil01, MPPU11, MS12c, MP01a, MP01b, MGH09, MNS06, MS06b, MV12b, Mit05, Mit06, MS08a, MKR07, MMK^{+}05, MSG06, MMK08, MMT08, MT99b, MTWD06, MR02, MKB^{+}09, MS00c, MPV01, MS08b, MVA01, Mon01, MMM05, MTU09, MSK^{+}12, MT05b, MD99, MR12c, MMC05, Mor06, MO05b, MBF08, MHW12, Mos01, MMH07, MSS04, MS10e, MR11, MSX10, MSS08, MAR05b, MB98, Mue01, Mue06, MCdS^{+}08, MC98, MCER10, MSUB04, MSKY10, MSR99, MdSC09, MV06, MA07c, MYS10, MWK^{+}09, MMM11, MNW98, MS11b, MV07, Mur99, MBH^{+}11, MZ06b, Nac01, NP02, Nad99, NL01, NWL06, NSM00, NW01, NNS03, NPPDG06, iNR08, NNPW12, NcC99, NR07, NYHO07].

Proceedings

[NMPM08, Nao04, Nao07, NR99, NdMO03, NO11, ND09, NJ10, NP12b, NZ06, NR98, NT06, NKPH08, NV08, Nes99, NSM07, NKLC06, NLKJ06, Ngo09, NKN^{+}05, NGHJ07, NJHJ08, NBGA08, NKC09, NP10, NL09, NCBT04, Nie99, NS00, NR00, NE02, NZ03, NKM^{+}09, NF07, NdG04, NL99, NWHR06, NV01a, NRWI09, NV01b, Nie03, NRR12, NBK04, NW10, NS98, NR04, NCJK08, NQL06, Nip98, NK99, NPS06, NP11b, Niv98, NT03, Nob12, NABK12, NBC^{+}06, Nor04, NSFB10, NR08, NQR09, NMP^{+}04, NBM09, NSL09, Nyb98, NSS03, O’B12, OSR^{+}02, OVG^{+}08, ONC^{+}11, OB07, ODP^{+}07, OT08, OT11, OS04, Oho03, OiDP98, OKHH^{+}08, OKS02, OAdGBP00, OYM07, Oka00, Oka04, OW07, OA07, OYY03, Oli00, OMMSP06, OS09, Oli07, OPS11, Omi00, OS03a, Ong05, Ong10, Ong11, OdQ12, OSC^{+}08]. Proceedings [OKdQ09, Oqu03, OWM04, Oqu07, OYA04, OSvL01, OT07, OWPY03, OALZ^{+}08, OD12a, OD12b, OKBMB10, ODV08, OR08, OV09, OST11, ÖRH12, OZ09, OZ11, PR08, Pag98, PRV11, PHR09, Pai09, PV07, PPP07, PS02a, PKM04, PKKG06, PHJ03a, PGM98, Pal03, Pal01, Pal00b, PS09a, PCH^{+}06, PAD^{+}06, PZE^{+}08, PCK^{+}12, PKM^{+}10, PDSD10, PDSN12, PP12a, PZK07, PZ08, POP09, PH11a, Par02, PA09, PR11, PBK^{+}10, PSST07, PV00, PCA^{+}09, PZYS12, Par09, PMÓ06, PSS^{+}10a, Păs09, PB07, Pat02, PdRM^{+}12, PC99, Pat03, Pat11, PMH09, PSE10, PFD^{+}10, PT08a, PAD08b, PDMFF12, PMT09, PMT10, PRK12, PMA^{+}07, PR05, Pel07, PV02, PLC11, Pel11, PH11b, PS02b, PH02, PCPS03, PD04, PF06, PF08, PF10, PFM12, PP99, Per01, PBH03, PR03b, PI05, Per07a, Per07b, Per08, PS08a]. Proceedings [Per09a, Per09b, Per10a, Per11a, Per11b, Per12a, Per12b, PT98, Per10b, PT08b, PCS^{+}07, Per11c, PS04b, PSD12, PW03, PVTG07, PSM10, PPS03, PMKG07, PY05b, Pez03, PHH10, PB09, PS03, Pfe06, Pfe07, Pfi00, Pfi01, PTM^{+}11, Pic01, PH12, PMDK00, PMWO02, PSNS99, POS00, PGD07, Pie08, PZ09, Pie10, PG01, PBK03, PK12, PH06, PP03, PT99, PPBL12, PRdRW09, PA03b, PB08b, PS11a, PRG09, PRG10, PRG11, PK00, PHOK10, PLG06, PP11, PMC06, PJ12, PVC^{+}99, Pol00, PS98b, PS09b, Pol99b, PC00, PBS08, PLK^{+}12, Por98, PR00, PS07b, PG09, PvS07, PISW05, PZ07b, PLI08, Pre00, Pre02, Pre09, PT11, PF07a, PWY08, PLMV12, PMS08, Pri03a, Pri06, PPM09, PM09, PRR05, PCA02, PPH07, PBW05, PS12c, PS11b, PW07, PW08, PSS98]. Proceedings [PPF^{+}00, PP04, Pup99, Qin10, QQ11, QOZ01, QIW07, QMW09, QSWL11, QBA01, QSBC98, QBY^{+}12, Qui98, QS00, Rab10, RF01, RCC12, Rah04, RF10, RiN12, RWSS12, Rai03, RG05, RWA06, RSV07, RMDS10, Raj09, Raj02, RFBW12, RW11, Ram01, RR08, Ram08, RRR99, RS10, RS06, RS09a, RR12, RA12, RM02, RAW07, RD01, RFZ03, RS00, RS05, RCJ^{+}08, RT12, RS99b, RO00a, RT07, Rau11, RMFG06, RRBE09, Rau04b, RLKHL12, RR98a, Rea01, RV10, RR01, RBG09, RD12a, RA00, RST09, RK06, Rei09, RW06, RRR01, REF^{+}06, Reu99, Reu01, RL06, RM04a, RS12, RK12, Ric02, RRM06, Ric98, RYG11, Rig08, RMTW11, RABR03, RF99a, RPBP06, Rob11, RR04b, Roc07, RL11b, RBsC^{+}06, RCLE05, RS09b, Rog11, RLT^{+}08, RK02, RKW04, Rol98, Rol99, Rol00, RV02, RV11]. Proceedings [RJS^{+}08, RWD09, REPH06, RS03, RT10, Ros03, RT09, RBGG06, RH98, RBC^{+}06, RS01, RV03, RSW12a, RO00b, RS09c, RD12b, RF99b, RJL^{+}08, RDK09, RG11, RMK07, RB00, RSK08, RGLA09, RBB^{+}06, Rus00, RZ12, RSTZ04, RTZŻ06, RTZZ08, RKS^{+}12c, RKNV09, RSK03, RS09d, RSW12b, Sab10, SBV^{+}11, SS07a, SPBP08, SP09, SNČ^{+}12a, SNS03, SN08, SNC12b, Sag05, SO07, SW07, SW04, SMD04, SKS02, SCH^{+}09a, SKS^{+}06, SS11a, SGSV10, SL11, SdSMO12, SGPP06, SL12, SS11b, SHJ09, SSO^{+}11, SYNC11, SW09a, SRG04, SYMA09, STP^{+}10, SK11, SZR^{+}08, SPCG07, San01, SRS03, SVS02, SKR11, Sar03, Sas05, SHM^{+}10, SNE08, ShK06, SMBQ07, SPH07, SHØ06, SSRA98, Sch04a, STT11, ST00, SH08, SS01, SDG^{+}07, SKMT04, SH12b, SUW02, Sch11, STL11, Sch04b, Sch06]. Proceedings [SCHS07, Sch09b, Sch12, SR07, SEG09, SF11, SB07, Sch01a, SMH^{+}12, SDR^{+}00, Sch10b, SW03a, SJW^{+}12, SvS00, ST12a, SW03b, Sch07, SS09b, SM06, SE10, SS06a, Sco02, Seb03, SLH04, SLZH07, SVS08, Sei07, Sei12, STT04, SP08, SG10, Sen10, SO09, SSJR10, Ser12, SC09a, Ses06, SVA05, Sex09, SEO^{+}09, SPK01, SMP07, SHX^{+}06, SW09b, ST02, SW08, Sha07, STS04, SWJX12, She01, SSD^{+}08, SS06b, SJ06, SKK08, SKSA12, SLC01, Shu07, Shu09, Shv06, SF08b, STA^{+}11, SCG98, SCR06, SKB10, SFN^{+}11, SCP^{+}12, SMK01, SMPU98, SEL09, SB12a, SMEGA10, Sir12, SL99a, Ska01, SLL01, SBH98, Slo99, SCH04c, Sma05, Sma08, Smi01, Smi12, Smo00, SC98, SG03, SLSW12, SPO11, SKP00, SKP02, SKP04, SKP06, SHKP08, SHKP10, SHKP12, SL99b, SAUL11, SZ00]. Proceedings [SK08, SBM11, Son00, SLLS03, SEN07, SPC^{+}08, SEN08, SMMG01, SQL10, SC09b, SMK02, SLKL09, SH07, SS08b, SMK06, SR11, SH09, SF08c, SXC05, SRY07, SB12b, SS06c, SSU12, SS99, SMCM07, SHL08, SvdMOP06, SDH03, ST12c, Ste98, SL03, SZ05, SH10, Ste01, SDK^{+}08, STTP02, Ste99, SWB03, SS09d, Sti90, SRM^{+}11, SSKR02, SCKR03, SRT06, SHS10, SD10, Sti94, ST01, SP12, STY^{+}07, SWMM06, SFV^{+}11, SGD09, SS10, SM01, Stu02, Stu08, ŠPTM00, SBH07, SBH09, Sub05, SHRM06, SLCT11, SNSR06, Sur02, SLM07, SMS12, SHYEF08, SHUH08, SWSY11, SC00, Swi99, SBH^{+}04, Syv02, SV12, SKR^{+}10, SH11, TR12, Tah00, Tah01, Tah09, TMLL09, TOOO07, TM09, TWQ^{+}12, TDWG11, TBE08, TMG10, TMdB11, TMS12, TSG00, TLF01, TFK11, TWYhL01, TLZ^{+}08, THL11, Tan11]. Proceedings [Tau08, TM99, TG04, TN03, TBKO10, TY11, TdLdOQ08, TYZ11, TTM10, TLC00, TNN^{+}01, TD07, TS10, TKCH09, TCYZ12, TW12, Tho99, TG00, Tho06, TW07, TJ08, Tho11, THX^{+}07, TG12, TBH03, TSS11, Tis02, TBJ02, TN09, Tiu00, TT06a, TQYX11, TLJ06, TBF^{+}09, TO08, Tok07, TNP^{+}07, TG02, TN04, TNM05, TNY07, TN08, TNI09, TND10, TNYL11, TEB09, TCJ10, TTF07, TBD12, TG10, TL02, Tra04a, Tra04b, TGAH11, TKHS05, Tre09, TZS00, Tri98, TMN09, Tro11, TDK^{+}10, TCLH10, TOL08, TdRM^{+}09, TSKGB04, Tuc01, THT03, Ued10, UPB07, UMN^{+}08, UBM02, Urz05, Uus06, UDF06, VvS99, Vad07, Vah09, VJKT00, VKV00, VGP08, VTK^{+}12, Val06, VJ11, Van02, VW07, VGM^{+}09, VM99, VM01, VW05, VV03, VWH06, VBH07, Vau98, Vau05]. Proceedings [Vau06, Vau08, Vck99, VJAA03, VA07, VG12, VC03, Ver08, VMBMN04, V^{+}06, VDD^{+}11, VKJ03, VZM^{+}10, VKMT11, VBS^{+}10, VMB^{+}12, VW08, VD12, Vit08, Vit10, VSH11, VZ99, VGB11, VS02, Vog01, VSTN06, VBCBS05, Vor02, Vor08, VLY09, VAM06, VP04, WF98, WHL09, WAS06, WRP07, Wag08, WSS04, Wal99a, Wal01, WKP03, Wal04b, WYL01, Wan01, Wan03, Wan04, WPV04, WJ05, Wan05a, WJS^{+}06, WPSY06, WLC^{+}06, WPRW06, WPR07, WPR08, WLGB^{+}08, WFZL09, WGMP09, WYSS10, WGLL10, WR11, WZD11, WLO^{+}11, WS12, WZJX12, WZH^{+}12, WB00, WT07, WZ03, WSN^{+}06, WSTI08, WY99, WZ09, Wat00, WY04, WK02, WR07, Wel01, WC98, WW02, WMR00, WLP^{+}09, WMS04, WL04, WHG07, WVD02, Wes01, Wha03, Whi00, Whi98, WCK11, WNE99, WTM^{+}02, WB09, WTP^{+}06, WP10, Wil01a, WR09, WM04, WM03]. Proceedings [Wim03, Wim04, WSGA06, WSG07, WSF08, WSJT09, WCJS10, WJP07, WFB12, WR03, Win11, Win09, WDKK07, WZ08, WHS01, WPD04, WZ11, WKS^{+}10, Wol07, Won01, WCBD08, WANL08, WKK98, WYX08, WLRL08, WD00, XCd03, XLKZ12, X^{+}12a, X^{+}12b, XYM^{+}07, XKGS12, XXX^{+}09, XBS^{+}10, XCR11, XL11, XZCL07, XYZU11, XBM12, Yah11, Yak00, Yak02, Yak04, YN06, YD98, Yam08, YEY04, YJ04, YJMU06, YJS^{+}06, YW06, YZC^{+}07, YDW08, YCC^{+}08, YYI^{+}09, YMHH10, Yan11b, Yan11a, YGE12, YLW^{+}07, YSP^{+}10, YRWS11, YT99, YLS^{+}12, Yas00, YŞ03, YI12, YKF^{+}06, YX04, Yi06, YAF^{+}02, YTC^{+}07, YWRS11, YCB12, YGGÖ05, YM01, Yor08, YHV12, YSR^{+}06, YKM06, YLLZ04, YKL06, YGL^{+}10, YLC^{+}10, YYH^{+}12, YY01, YZCW07, Yun02, YDKM06, ZRP^{+}09, ZK06, Zam08, Zan00, ZTTY12, ZBRL12, ZdC08, ZCCL10, ŽL08, ZGK^{+}07, ZCRL08, ZPT^{+}06, ZL98, ZS00, ZJ04a, ZZ04a, ZJ04b, ZTY^{+}05]. Proceedings [ZS07, ZOCJ07, Zha07, ZYBX08, ZPTI09, ZZP^{+}10, ZS10, ZO10, ZHLW12, ZG03, Zhe02, ZJL06, ZZ99, ZSO99, Zho01, ZRTS03, ZLLC09, ZCGH11, ZYH03, ZJXC03, ZNCF03, ZZO03, ZSP^{+}04, ZOM05, ZYB06, ZSY^{+}06, ZLY07b, ZZTG07, ZW09, ZYDL09, ZY10, ZZLL10, ZYRH06, ZF05, ZBH12, ZKZ00, ZJS^{+}02, ZT04, ZVU99, Zit01, ZBY07, ZBG98, Zuc03, Zui99, ZQ98, ŻR99, dA09, dBBC02, dCVK07, dFEN04, dLZ12, dLWC03, dMB05, dMS06, dMLG03, dMS09, dO98, dRF01, dRS01, dRWK^{+}10, dSK12, dSDL04, dSTP11, de 98, dS11a, dVGR04, dCHM12, dlBP08, tTVH^{+}12, vBC08, vDLJ11, vEGW09, vEGSW11, vWN99, vG02, vHV08, vHC08, vHH09, vLWH^{+}00, vLIvdH^{+}07, vLMP^{+}10, vO04, vSN01, vSH06, vdPW10, vdBHOS09, vdBV01, vdHI99, vdHXMW08, vdAB03, vdAHW03, vdL98, vdL01, vdMBD08, vdMvdT08, vdPS08]. Proceedings [vzGIK08, AI97, AB10, AP93, AS94, AGK^{+}10, AN97, AH94, AK97, AF97, ABLM05, Akl95, ALS^{+}05, ASP01, AN95, AM91c, AJM87, APO96, AW90, Alo05, AH96, AS06, ABD04, Ame91b, ABS91, ADW05, And94, And96, And95, Ang97, APS06, ABG96, Apo92, Apo93, AH97a, AM04, ACP05, AGS91, ACJT07, AJ94, Asa90, Asa96, AB81, AP83, ADCR89, Aus93, ADL09, Aya95, AB97, BV05a, BSMM04, BM96, BPSP06, BL05a, Bak95, BC95, BL92, BTT90, Ban92, Ban93, Ban94, BBP89, BSW95a, BGLT05, BCKK05, BMSMS04, BG93a, BMH05, BHP95, BL04a, BG91b, BBKP06, BEM91, Bec05, BIRT05, BB95, BC92, BR92, BD97a, BCT06, BV92, BCD05, BdR94, BS92, BMP90, BR89, BGD^{+}05, BU94, Ber96, BR97, BC88, BB89, BG93b, Bha95, BN91a]. proceedings [BCB97, Bih97, Bil97, BR96a, BMW93, BH92, BN91b, BFP04, Bjo80, Bjo87, BHL90, BBP93, BBP96, Bla05, BTPPS05, Bla88, Bla90, BHR90, BMJ88, BMM07, Bod91, BRW93, Bod96, BL89, BS94, Bv97, BC07, BBD97, BCP94, Boo91, BF89, Bor91, Bor92, BK96, Bor94, BP05, BS93, BAL88, Bos96, BM94, BMYZ91, BMYZ95, BCRT92, BP04b, Bow94, Boy95, BG05, BDV95, Bra96, Bra90, Bra93, BCH^{+}05, BJS93, BW05b, Bri97, BL05c, Bro92, Bro94, BW92, BP93, BRR97, BDW97, BV94, BC04b, Buc90, BBL87, Bud91, Bun94, BS06c, BW97, BMP96, BFKO05, yCW96, yCCZ07, CIM^{+}05, Cal86, CC93, CCP96, CL96, CDP^{+}05, CE95, CC05, CNX05, CFB^{+}05, CS11, CO94, CTT93, Cha05a, CV96, CC91, CP88, CJRT05, CPGL05, CL04, Chi95, CC97a, CJ05]. proceedings [Cho88, CCLR98, Chr91, Chu03, CNY95, CH96, CD05, CK91, CKM93, CCK95, CG88, CW89, Coh92, CMM93, Coh94, CGM95, Coh96, CM05a, CH91, Com97, CFR92, Con97, CN05, CMV96, CLT05, Cop95, CLP05, CS97, CFP94, CC97b, CP05, Cou93a, CCR86, Cou93b, CS96, Cra05, CG94, CK93, CDG89, CM05b, CMQ96, DKP05, Dal92, DWSK05, DH05, Dam91, DLV04, DD97, Dar97, DK90, DSM94, Dav91, DG96, DV05, DN04, DR05, DDFM05, DH97, DD95, De 95, DE05, DGMS97, DSS89, DFK91, DSS91, Deh93, Deh97, DLOS05, Dei90, Del97a, Del97b, DKM91, DDJ00, DL04a, DD06, DY06b, Den85, DJ90, DM90, Der92, Der89, DL95a, Des94, DWML05, DEQ92, DCM82, DCP95, Di 97, DKD05, DH94, DR81, DS96, DJRZ06, DF96]. proceedings [DD05, Dil94, Din04, DJNR09, DKU05, Dit88, DJS91, DFN97, DH06, Dji89, DKB03, DGR92, DW94, DMW96, DBB^{+}04, DvV96, Dos93, DER94, DZ94, DL95b, DKOZ05b, DKOZ05a, DL81, DG97, DNS95, Dyc94, DHSH96, EV96, EHP94, EH95, EKR91, Ehr91, Ehr04, EEM96, Ell95, EKT94, EG97, EW91, EFW93, EHSH92, Esi93, ER05, EKL^{+}05, FS05, FKP^{+}05, Fau90, FM97, Fei91, FR95a, FR96, FR97, FSM03, FR95b, Fer96, FHRR05, FS09a, Fil91, FD93, FO95, Fit93, FNPW05, Fra05b, FP05, FC93, FK95a, FGM92, FGL96, Fre96, FT94, Fri94, FK95b, FW92, FM91, FLS05, FvdHJT10, FG95, Fum97, FG04b, Fur05, FM96, GO94, GO96, Gab97, GG05a, GP05, GU95, Gal97a, GHM05, Gal97b, GMV05, Gan96, GL97, GM01, Gau93, GHR89, GH94a, GH94b]. proceedings [Gia04, Gia89, Gib89, GHK97, GL05, GK95, God05a, God05b, Gol94, Gol96, GLIC10, GPPW08, Gom01, GS95, GKS05, GS88, GN90, GLM93, GV95, GLM97, GLW88, Gra05, GL92, GT93, GP89, GSV05, Gru97, GW93, GNR94, Gue90, GQ95, GS94, GF97, Gün88a, GS91, Gun94, Gut94, Gyi96, HMRT05, HM94, HM08, HI93, HT97, Ham89, Ham91, HOQ97, HS05a, HR96, HHM97, H^{+}05, HB97, HK94, HP95, HS94, HG96, HK92, Heb95, HL04, HCS^{+}05, HR95, Hel94, HT05, HP94, HS95a, Her92b, Her91, HS95b, HS97, HML97, HM96, Hir97, HF97, HS95c, HSS96, HK05c, HGZ05, Hof01, HK96, HPP88, HKNW07, Hsi95, HM04, HSB^{+}96, Hug91, HP89, HW05, HM05b, Hut94, Hut96, Iba92, Ibr95, IRM93, IKY05, Ish97, IH90, IM91]. proceedings [IY95, IHQ96, JPBT05, Jac04, JY96, JW05, Jan88, Jan87, Jan89, Jan92, Jan93, JSZ95, JBJ94, JS97, JKB04, JS05, JKB96, Jen92, JE06, JL97, JRJ05, Joh97, JP04b, JP96, JS84, Jou94, JS94, JK05a, KMP84, Kah87, Kal97, KC05, KL95, KK97, KMR^{+}05, KJ88, KO91, Kap92, Kar91, Kar94, KL96, KKM04, KN91, KP92, Kat92, KL05, KJ91, Ker97, KD05, Kil05, KM96, KW90, KL92, Kir93, KKKN05, Kit88, KD99, KS93, KN83, Kob96, Kod91, KZ97, Kot86, KKD04, Kra98, KMS89, KSH05, KS91, KS96, KMN94, Kui92, KK95, KW96, KA05, LY04, LM93, LdRV94, LS92, LB05, LW95, LD97, Le 94, LP97, LG94, LYMM05, LB90, Len93, LGG97, LIJ97, LLC^{+}05, LS88, Les87, LA94, LM97, Li98, LST^{+}05, LWD05, LV05, LCHS96]. proceedings [LSS^{+}04, Lin14, LMV95, LKC93, LG95, LS89, LR94, LCB97, LSL04, Lom97, Lom93, Lon90, LQO04, LFL05, Lou92, Lou94, LG05, LCG97, Luk97, Luo05, LKZCS05, LO88, MPP94, Mad89, MMD05, Mah85, Mai88, Mai90, Mal97a, MW91, Mal95, Mal97b, Mal05, MPF95, MAD^{+}04, MMR07, MNT95, MS96a, MSH85, MR97a, MSZ90, MST92, MPB05, MD04, Mar93, MBS05a, MR91, MJR04, MMT97, MMP05, MMR91, Mau89, Mau91, Mau96, MT97, MSW05, May93, MST95, McC96, McC97, MH05, MS96b, MT05a, MTH05, MXZ05, MC94, Mel86, MV91, Men07, MT89, MMR93, MM96, Mic90, MD96, MMMO96, MMU96, MHK05, MP93, M^{+}06, MMS93, Mio90, Mio93, MCP93, MS95, MMMD97, MT86, MR97b, Moh91, Moh97, Mol95, MAA91, MR95, Mon96, ML95, Mor89a, MM97, Mor95, MK96]. proceedings [Mos97, MSH05, MB87, MWJ97, MBC^{+}97, MZF94, MB97, Myc95, MO92b, MC05, Nag90, Nag95, NDFMS97, Nak92, Neh88, NT92, NM94, Nes95, Ng93, NW97, Nie05, Nik05, NY93, NS81, Nor86, Nor87, NK88, Nor90, NS90, Nor97, NU92, eOP05, OP97, OA93, Ott87, OPKR05, PBB05, PS04a, PVV05, PCG^{+}05a, PGA^{+}05, PS97a, PS98a, PFCS05, Par85, PF05, Pat90, PW92a, PW92b, PZ96, PT05, PH97, PS96a, PpWR96, Per06, PT92, PM96, PT97, PVP05, Pet92, PN93, Pfe94, PMD90, PMD92, PMD94, PMDA96, PMD97, PSN95, PS96b, PV04, Pin89, Pin95, PDG92, Pit86, PPR87, Pit89, Pit91, PRJ95, Pla96, PB97, Poi06, Pol86, Pol97, PZH96, Pre95, Pri91, PRR94, PR97, Pro96, PB92, QMLW05, QV90, RCB^{+}04, Ral94, RCA90, RS97a, Ran93, RVY05]. proceedings [RZ91, RK93, RZ94, RM96, RS97b, RMS04, RCT05, Rei95, Rei89, Reu97, RMS^{+}05, RT95, RAM05, RB91, RV05, Rol97, RBC93, RBS93, RKM06, RSvdP05, RPZ97, Rov90, Rue93, Sah13, Sak91, SKMR94, San92, STC04, SL05, San94, Sat97, SB95, Sch95, Sch93b, SSW92, SCM88, SS91, SS94, SE94, SV97, SS05b, SPZ90, SH91, SWAF95, Sch92, SM91, SLH05, SP90, SZ93, SZ92, Seh97, Sel95, SZC^{+}03, SOS96, Sha86, SA96, SLL^{+}06, SdDK91, SILS05, Sho84, Sho05, Shy92, Shy93, SU97, Sie86, Sif90, SJT05, SRW98, SSAP05a, SSAP05b, Sko85, Sle92, SBG04, Sma97, SF96, Smo97, SPSV^{+}04, SK97, SDP97, SP93, SLPM96, Spi95, SLF91, SC96, Sta95, SA97, SSv94, Ste95, SSB90, Ste94, SW97, SLP05, SKB98, Str99, Str96, SDM05]. proceedings [SV05, SA04, SC05, SKN05, TT95, TTP05, Tar91, TC95, TC99, TV94, TDS94, TDG91, Tha96, Thi94, Thi95, TS87, Tis94, TT05, TWN92, TP94, Tom91, Tom92, Tor93, TSK92, Tou94, Tou96, TV87, TGM97, Val94, Van97a, Van89, VP96, VPM97, VA95, VDPD96, VH04, Vit95, Vog88, Vol93, Vor92a, Vor92b, Vor93, VWY97, Vyt91, WT96, WC95, Wal04a, Wal84, WTZ^{+}04, WCO05a, WCO05b, WCO05c, Wan05b, W^{+}12, War94a, Waś96, Wat95, WS96, Wet92, WH95, Wie99, WGL96, WBW94, WTGA05, WGD05, WN96, WLS89, Wol95, WJ95, WMT96, YZZ^{+}05, YAL^{+}05, YRMD05, YBL^{+}04, Yi05, YI91, ZBW94, ZEL04, ZL96, ZHF04, ZJ05, ZLJ05, ZJ06, ZGT05, ZLDB05, ZLS^{+}06, Zil87, Zim92, Zip92, ZS05b, dBdRR90, dBdRR91, dB92, dBdRR93, dBdRR94, dSG05, ter97]. proceedings [vB05, vE91, vL88, vL89, vLS91, vL94a, vL94b, vW97, vdABCC05, vP93, vM94, vGH96, von96]. Proceedings. [CMM09b, DA08, HSMPA09, WZA10]. Proceedings/ICICS [QGZ03]. proceeedings [GGHN97, Ish99b]. Process [ABMM07, ABD^{+}09, ADR07, BVO10, BGK12, BI04, BKS05, BH06d, COBV11, CS05a, CGBDA08, DEKR09, DKW99, DPW04, DJJ12, DRS08, DBB^{+}08c, DFS06, ED06a, HMT10c, JS08, JvdA09, KF74, MV06, MA07c, MYS10, NM08, OKS02, Oqu03, Rei03, RRM06, RMTW11, WPRW06, WPR07, WPR08, WGMP09, tHBP08, vdADO00, Amb01, BO00b, BKS01, Bra76, Con00, Der92, Din04, Gru98, LBO05, Mon96, RAM05, Sch95, Sch96, Sch98a, War94a, vdAHW03, vdABCC05, DG01, HS02]. Process-Aware [JvdA09]. Processes [AMS06, BG09, Bra80, BKZ05, CDSS07, EB04, FLMG10, FKM08, MMM10, MP01a, MYS10, Rei03, RtTMP10, WGMP09, AJS05, BMH05, Cas01, CN98, GF12b, Gue90, Hai82, MS03a, SW75, Zwi89, MvOvRdV05, PPdV12, Wes99, Bur97]. Processing [AJNJ12, ANS04a, ANS04b, ANS04c, ADB^{+}12a, ADB^{+}12b, ATM09, ABJ02, Bar02, BEM91, BMYZ91, BMYZ95, BFMR96a, BFMR96b, BFT^{+}08, BIMW12, BZ08a, Bur90, CBGCCZ12, CPRS01, CVTP12, CAS^{+}09b, CL10, CJ08b, CHG^{+}07, CS05b, CMOZM11, CFS06, DGT10a, DGT10b, DBM09, DGJ09, ELNM08, ELN^{+}10, EML^{+}12, EVV^{+}11, FZJE^{+}05, FZJE^{+}06, FR00, FRS03, FSV09, FS07d, FS08, FS09b, FS10, FMPT10, FHW98, GMS03, Gel04, Gel06, Gel07, Gel08, Gel09, Gel10, Gel11b, Gel11a, Gel12a, Gel12b, GG98, GSS11, GVT^{+}11, GYD^{+}06, HHJ^{+}86, HDBM11, HK05b, HGZ05, HRM^{+}10, HMMW10, HYPY10a, HYPY10b, HC09, HXC^{+}08, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e, HKH10, HMCL06, IAL^{+}07, IK12, IDMY08a, IDMY08b, JNR11a, JNR11b, JMC10, JRPP07, KPS12, KP06, KTH09, KL07b, KLM^{+}07, KBP07, KS06]. Processing [KM04, KWCW06a, KWCW06b, KWCW06c, KFMM06, KKC09a, KKC09b, LO02, LMSS07, LLC09a, LLC09b, LPZ12, LMA09, LAF^{+}10, LXH^{+}12, LRH10, LST^{+}12, LZK11a, LZK11b, LZK11c, LMB08, MF11a, MF11b, MBT^{+}03, MM09a, MSWZ06, MM04, MMM05, MSX10, MWK^{+}09, MMM11, NWL06, NdMO03, NJ10, NR08, PKM04, PBK^{+}10, PPM09, QLK^{+}10, QLK^{+}11, RM02, SGPP06, SEL09, SPO11, SCZ10, STY^{+}07, SC00, SH11, TN03, TY11, TdLdOQ08, TYZ11, TGAH11, VY93, VMBMN04, V^{+}06, WR09, WLS89, WPZA11, WMB10b, WMB10a, WDPW02, WDPW04, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, XCHZ11b, XCHZ11a, XSA^{+}12a, XSA^{+}12b, XZCL07, YBM87, YJKK06, YJKSW10, YAKM^{+}12, ZJL06, ZJXC03, ZYRH06, ZBY07, ABD^{+}99, AM04, AGL06, BG93a, BE87, BLKW01, BK81, BCRT92, BKM01, BDV95, BV94, CYGL04, CNX05, CCH02, CPGL05, CSV05]. processing [Chr00b, CH91, CKMS05, CN05, CM05b, DWSK05, DLV04, Del97a, Del97b, DP97, DL81, DG97, FR95a, FR96, FR97, FR98, FR01a, FRS02, FRS05, FFRS05, yF75, Fil91, Gel01, Gel02, Gel03, Gel05, GY03, HAM95, HK05c, IL01, Jah93, Kay03, Kli85, KBH03, KSTP99, Kum90, LGG97, MML97, Mey95, MF02, MNW98, NW97, NK99, Ooi90, PDH99, PDH01, PCG^{+}05a, PW92b, PVC^{+}99, PR98b, RR89, RV05, Rol98, Rol99, Rol00, RF99b, Sak01, SG96, Stu85, STLK05, TPB79, ZG03, ZVU99, vM94, Jae01, RB91, SFZE07, Fra96, Len90, Mit98, RS96, SLC01, Wal82]. processing-PCM [SLC01]. Processing-Systems [DP97]. processor [DTV02, FKS86]. Processors [Pfl02]. ProCoS [LdRV94]. Product [ABD^{+}09, BL10b, Cha02b, CGBDA08, DBB^{+}08c, Nor04, vdL04, eOP05, SLF91, vdL98, vdL01, vdL02, BVO10, BO00b, BKS01, BI04, BKS05, COBV11, DJJ12, JS08, MV06, MA07c, OKS02]. Product-Family [vdL04, vdL02]. Product-Focused [BVO10, COBV11, DJJ12, JS08, MV06, MA07c]. production [Ish94, LL95, MM82]. Productivity [Sch11, OZ01]. PROFES [BVO10, BKS05, COBV11, DJJ12, JS08, MV06, MA07c, BO00b, BKS01, BI04, OKS02]. Professional [AH97b, ADB^{+}05]. Professor [MFÁ^{+}09b]. Profiles [BHMS07, GR06b, ZSSM03]. profiling [MS00a]. Prognosis [MDY^{+}10]. Program [Alp11, APM08, BEH^{+}85, BM02a, BDK10b, Bru04, BL04b, CDE^{+}07, CS12b, De 10b, GN12c, Han09, Hil06, Kin08, Kno98, Koz04, Pet02, Pue07, Tah00, Uus06, Vid12, BO00a, BCG02, BMW93, Bos00b, EDD04, Eng74, Eta05, FS03, Fle99, FT94, Fuc98, Gal97a, Gre75, Jeu98, Lau01, LBCO04, Leu03, MPS93, Mol95, Pet80, Pro96, Roy87, Tah01, Van89, BB79, BL04b, HKB93, KMS06, LK09a, RSB07, Sta82, Ley01]. Programmable [BPV04, GH93, HS94, HG96, HDLM09, LIH99, VT97, CCdS03, HG00, BW01, GZR02, GH93, HS94, HG96, HK98a, LCG97, ML95]. programmation [Rob74, Rob80]. programmed [Tha84]. Programmers [AH97b]. programmes [GB77]. Programmiersprachen [SF74]. Programming [AM91b, AS11a, ADG^{+}08b, AHS11a, AHS11b, AB11, AMSM^{+}09b, ABDW10, Ame91a, ABH^{+}10, AFR^{+}04, AS07b, Arn90, ABFR07, BV02, BG03a, BGHJ07, BS11b, BFGM05, BGLT05, BBS07, Bar11, BGP11, BCT02b, BS06a, BJP12, Ben06, BCK08, Bes07, BHCL87, BN04, BV12, BRST08, BKV10, BAL88, BS03b, BJ01c, BPSW06a, BPSW06b, But10, CKS04, CNP^{+}08, Cas09a, CVV08, CZG^{+}08, CV10, Cod01, Coh10, CTE^{+}06, CDSS07, CM00, CS02b, CMPW12a, CMPW12b, D’H10b, DN07, DN88, De 07b, DFKM08, De 10a, DS09b, Dec00, Deg03, DF11, DV07b, DCM82, DO89a, DO89b, DKLS04, DP10b, Dro08, Dro09, DS06, EOE^{+}07, EB04, Ehr85b, EEK^{+}87, EKLM87, ES10, ELS09, Ern07, EN74, Esc10, EAES^{+}10, ET06, FRS08, FPFR05, FV10, FW07, FZ86, FL11, Fre98, FTF91, Gan88]. Programming [GG09, GH08, Gen09, Gib12, GN88, Gor10, GW11b, HW06a, HS74, Han81, Han98, HV06, HW09, HCF06, HY03, Hor06, HPSZ08, HPZ10, HM11c, Jon90, KS02a, KS02b, KS04a, KV04, KOL^{+}04, KPS09, KP05, KB92, Kuc11, LO11, LS04a, Le 02, Lee11, LN03a, LS06a, LPR08, LMT10, Mad92, MP98, Mar11c, MW03, Mez11, MSK^{+}12, MCdS^{+}08, MOTN07, MTNL12, Nie96, NV01b, NJ09, NJA^{+}11, Nob12, NABK12, OVG^{+}08, O’S03b, OFR09, Ode04, PHZ11, Pal03, PV00, PR84, PP12b, PT08b, PS03, Pri03b, Qin10, Rao92, RR04a, Rob80, RW86, Ros03, Sag05, SM11a, San01, San94, Sar03, SF74, SKP10, Sch04b, ST12a, Sch02, Sei12, SHW09, Ses06, Sha86, SFN^{+}11, Smo00, ST12c, Stu02, Stu08, Swi99, Tho06, Uni81, UA83, UWB^{+}06, Van02, VW07, VGM^{+}09]. Programming [VV03, Vit08, Wal01, WW02, WTM^{+}02, ŽL08, ZHP12, dlBP08, vHH09, AG01, AS94, AMS06, AGK^{+}10, ABL96, ADG^{+}08a, ADR01, AM97, AMSM^{+}09a, Ang97, ACJT07, AB81, AJ01a, AJ01b, AB78, AP83, ADCR89, BV05a, BLPW03, BC89, BCB99, BC95, BL92, Ban98, BM05b, BMH05, Ber00, BK05b, BBRM98, BBP93, Bla05, Bor94, Bra85, Bre91, Bro92, Bro94, BW91, BW92, BP93, CIM^{+}05, Car03, CS11, Cha05b, CH98, Coi96, CFR92, CDW86, CJE82, CBW99, CCR86, CMQ96, CF00b, DGMS97, DL04a, DM90, DHM00, DR81, Dia83, DPP95, DFN97, DPP97, DPP98, Don76, Dyc94, DHSH96, DF99, EV01, EFT01, EHSH92, EK81b, FKS86, Fos02, Fre96, FT94, FG95, FTF88, FTF89, GG05a, GVM99, GLP99, GG02a, GHK97, GL05, GT87, GLW88, Gre94, GNR94, Gue99]. programming [GW00, HR96, HHM97, HP95, HP94, HRA02, Hug91, HR99b, IY95, JY96, Jaf99, JLM00, JM95, JJ03, JM78, Jou85a, Jul98, JK05b, Kah87, KTC^{+}05, KW90, KL92, Kir96, Knu01, Kot86, KSW90, KU01, Kui92, LM93, LSW98, Lau93a, LMS96, LD97, LS88, LA94, LKC93, Loe74, Mag02, Mai88, Mai90, MW91, MS03a, MNT95, Mar80, MS03b, Mau79, Mel86, MM96, MS99b, Mil12, Mil01, MAA91, MR95, Mos97, Mue01, Mug97, Nad99, NY83, NS82, Nie93, Olt95, OSvL01, Ott87, Pag98, PGM98, Par84, Pat90, PD96, Pfe94, Pit86, Pod95, Pol99a, Pol00, Por98, RS01, RSK03, SS77, SH91, SHGB05, Sip97, Smo97, ST87, SV05, SHO99, Tis94, Tob80, TP94, Van05, VU05, VJ99, Vor92a, Vor92b, Vor93, Vos03, Wad86, Wad87, Wal04a]. programming [WMR00, WF77, ZEL04, vB05, vWN99, vLdB80, Bar11, BJ84, Bis04, BDDS05, BDDS06, BDDS07, BBT10, BW92, BP93, Cas09a, Chi04, CDN12, DDE04, DSRW08, De 07b, Deg03, DBB12a, DLM89, DM90, Dro08, GHK97, Gor10, Gut94, Hac75, Han98, HP94, HS95a, HPS09, Hu09, dCJB12, Kob06, KS96, Le 02, MW91, Oho03, Ram08, Rob74, Sag05, San01, San94, Sch04b, Sch02, Sei12, Ses06, Sha07, Smo00, Swi99, Ued10, Yan11b, Yi05, AP96, Gol82, Hol89, Jac91, NCW97]. programmming [Pet92]. Programs [AM07, ACÇ^{+}12, BCD04, BDd09, CDNS00, GB77, HM12b, JS11b, Koz82, KC84, MSH08, Sch03b, ANR99, BN94, BG91b, BC96a, CLMP02, DNS95, Eng81, FPMW06, GW03, GPM98, GHK97, HS95a, KS96, Luc87, Mei89, Mul93, Par85, Pet80, Sal83, Tau89, DF01a, FLMG10, GJ86, MO06, Mül02, Bar85, Fre87, Plu90, TU94]. progress [BJK06, AB10, AMGJ12, AP11, AS02b, BA99, BL06a, BCE09, BCD05, BL10a, BC11a, BC10b, BJ01c, CV04b, CRD08, Coe98, CC97b, DV05, FD93, GN12a, GG10, HN12, IFL^{+}12, JM03, LLMR09, MMV06, MPF95, MTOK06, MS02b, MV12b, NSM07, Ngu06, NP11b, PA03b, Pre09, RCC12, RAW07, RD01, RO00b, RS09c, RMK07, RSK08, SK11, SRS03, STC04, SL05, SPZ90, SRY07, SFZE07, Vau08, Wat95, Yao95, ZYBX08]. ProHealth [tHBP08]. Project [BBB^{+}85, BEH^{+}85, DFT04, DFF06, LT11a, MS99a, WH11, NM08, AS02b, BW91, HR91]. Projective [Heu04, Her92a]. Projects [JCC^{+}12, Mül07b]. Prolog [Bar03, Bee89, Nai86, HR99b]. PROMAS [BDDS05, BDDS06, BDDS07, BBT10, CDN12, DSRW08, DBB12a, HPS09, DDE04]. Promises [ÁCD^{+}12, DGG^{+}11b]. PROMS [BMO02, vSN01]. Proof [ACG^{+}07, CLKK07, MPS98, NPW02, GLM93, GLM97, Zwi89, KSHS01]. Proof-theoretic [ACG^{+}07]. Proofs [AM07, BH08, BCD04, BDd09, BJ12b, CDNS00, Dub09, FLMG10, FG10, GW11a, GM07b, HM12b, JS11b, MSH08, ANR99, BL05a, BN94, BC96a, CLMP02, DNS95, FPMW06, GW03, GPM98, Plu90, Sud95]. PROPER [CAS^{+}09b, GVT^{+}11, LAF^{+}10, ADB^{+}12a, ADB^{+}12b]. properties [BJM^{+}05, BRR87b, Pin89]. Property [Gol11a]. PROPOR [CVTP12, MBT^{+}03, PBK^{+}10, TdLdOQ08, V^{+}06]. Proposed [Uni81]. ProSim [WPRW06]. Prosody [Kom97]. PROSPECTRA [HKB93]. prospects [AAN00, DK90, MT99c]. Prostate [MDY^{+}10, MDHB11]. Protaras [DKB03]. Protecting [GKL09, GKL09]. Protection [CW94b, LM03, LSW12]. Protein [GI03]. Proteomics [IB07]. Protocol [AL11b, CZZ08, DV09]. Protocols [BMO02, BPS06, BAP^{+}05, CCMR02, CCMR03, CCMR05, CCMR06, CCMR07, CCMR09, CCMR10, CMMR11, CCMS11, CMSA12, GCC^{+}02b, Lin03, MKR^{+}04, RR04b, VC03, CCLR98, Chr99, Chr00a, Chr01, Lom97, BMO02, CZ02, vSN01]. prototype [LCS91]. prototyping [GH93]. Provable [BBCL08, BC11b, HK10a, PZ09, SLM07, TWQ^{+}12]. Provably [JN01, LdRV94]. Provenance [FKM08, GF12b, MMM10, MF06, FKM08, GF12b, MMM10, MF06]. Provence [Oli07, BTQ02, HCT02, KWA02, TL02, Vää99]. prover [Pau94]. Provers [Wie06, KK95]. Providence [Bes07, DST01, HKS09, Ish11, LP97, Sil01, SRW98]. Provincetown [McG08]. Proving [BW03b, BNUW09, BF12, CMT02, KP10a, MMT08, SB07, vEGSW11, AH00, BHP95, BDH^{+}99, BJ01b, GN98, GF97, HM05b, JS94, MC94, MMMO96, SWAF95, SBG04, ST87, vGH96, San80b]. Provisioning [SF10, SSKR02]. ProvSec [BBCL08, BC11b, HK10a, PZ09, SLM07, TWQ^{+}12]. Prozessrechner [KF74]. PSD [DFF06, DFS08, DFM11, DFT12, DFT04]. PSDML [DGDM^{+}11]. PSI [BBZ00, CVV12, PVV10, VV07, BBZ01, BZ03]. PSIVT [CL06a, Ho12a, Ho12b, MR07a, WHL09]. PSL [ST12b]. PSLS’95 [IHQ96]. PSSE [BCSW10, CSW06]. Psychological [PBSB09]. Psychology [Har07, Har09, Har11a]. PTCS [KSHS01]. Public [BDZ04, CFGN11, Cra08, Des02, FBM12, IZ98, IZ99, IZ00, JT09, Kim01a, NP02, NP10, OW07, Vau05, YDKM06, AL06, BDZ04, BFS92, CL11, CFGN11, CZ05, Cra08, Des02, FBM12, IZ98, IZ99, IZ00, JT09, KGL04, Kim01a, LSF07, MP10b, MMK08, NP02, NP10, OW07, PNPP12, Vau05, YDKM06]. Public-Key [Cra08, OW07, YDKM06, BFS92]. publication [HNR05]. Publishing [KM04, HAB98]. Pucón [SK11]. Puebla [AB10, BS11d, BS11e, LRG04, MTCOK10, STC04]. Puerto [CMM93, PCPS03, Rol99, RF99b]. Pune [Jos00, Nor87, NK88]. Punta [GPV00]. PVM [CHD07, DKD05, KKD04, LKD08, MTWD06, RWD09, AD98, Bod96, BDW97, CD01, DLM99, DKP00, DLO03, KKDV02]. PVM/MPI [CHD07, DKD05, KKD04, LKD08, MTWD06, RWD09, AD98, CD01, DKP00, KKDV02, BDW97, DLM99, DLO03]. PWC [COB06, CGGO03, NdG04].

Qatar [BJM09, Cer07, CVV08, HZLL12a, HZLL12b, HZLL12c, HZLL12d, HZLL12e]. QC [BBD^{+}10, ELN^{+}10, FGK^{+}10, JP11a, JY14, Mil12, Smi12, SRL12c]. QCQC [Wil99]. QCQS [Wil99]. QI [BSL^{+}09, BDLMM12, SMF^{+}11]. Qingdao [CGL^{+}11, HHL07b, HHL07a]. Qinhuangdao [LC11]. Qld [ZL98, ZL97]. QLQP [GHI^{+}06]. QofIS [SPSV^{+}04, SSKR02, CRS00, KS03b, Smi01]. QoIS [ALS^{+}05, HRK^{+}07, HP09, RBsC^{+}06]. QOS [ABLM05, CMS11, SPSV^{+}04, SSKR02, SRT06, RST09, SCKR03, BKB03, MB01, Mar03]. QOS-IP [ABLM05, MB01, Mar03]. QoSA [BPR08, HKP10, HCR06, MGH09, OSRS07, RMS^{+}05]. QSI [MTH08]. QuaCon [RFBD09]. Qualitative [Gab97, Gab97, Her94, Ren02]. Quality [BPR08, BF11, CPV03, GH09, Han02, HKP10, HCR06, HS11, KS03b, KC02, MGH09, OYM07, OSRS07, PR08, RD12a, SPH07, SPSV^{+}04, SSKR02, Sto04, WP10, AH97b, BKB03, Nes95, RMS^{+}05, ABLM05, BPR08, CRS00, Her02, HCR06, JSW03, KS03b, MB01, Mar03, Nau02, RMS^{+}05, RFBD09, Smi01, Smi03, WHS01, dMB05]. Quality-Driven [Nau02]. Quantified [Her02]. Quantifiers [Kai11, Vää99, Bur91]. Quantitative [ABDW10, Liu11, Mel07, SC09b, BB01a, CKM93, FK95b, Gab97, God05b, HP99c, KS91, NZ03, BB95]. Quantum [CM09a, KM08b, vDKS11, Wil99, BSL^{+}09, BD08, BDLMM12, Sen10, SMF^{+}11, Wil99, Yan11a, Gab13]. Quebec [CFG96, Ham00, Kro00, LM06a, MMS93, BDK10b, Mil12, PU04]. Queensland [DG96, McC97, SZ93]. Queenstown [KKS11c, KKS11d, KKS11e, KKS11f, KH11a, KH11b, Zhe02]. queries [Fre87, Has96, MLK04]. Query [AMCL02, AYB^{+}09, CHAL04, CDY^{+}11, EK12, Jae01, LPOA^{+}06, Nau02, ACL98, Ooi90]. Querying [Aso04]. Quest [Her02]. questioning [Ley01]. Queueing [Dad01]. queuing [LT01]. Quisquater [Nac12].

R [BDFM12, GG98]. R1040 [BP95]. RACE [BP95]. Radical [HRR^{+}06, THR03, WKB04]. Radio [Yal10]. Raesfeld [HHJZ12]. RAID [BSC12, JSK10, KJB09, KLC07, LKT08, SBM11, VZ06, ZK06, DMW00, JVA04, LMW01, VKJ03, WVD02]. Railway [GKS^{+}07]. Rainbow [CRS11]. Raleigh [NQL06, ZCRL08]. Ramat [HBN06]. Ramat-Gan [HBN06]. Ramatuelle [Pin89]. RAMICS [dS11a, KG12]. Randell [JL11]. RANDOM [CJRR07, CJRT05, DJRZ06, DJNR09, GJRR08, GJRR11, GJRS12, SSJR10, AJRS03, GJRT01, HJRS99, JKRR04, RV02]. RANDOM-APPROX [HJRS99]. RANDOM’97 [Rol97]. RANDOM’98 [LRS98]. Randomization [AJRS03, CJRR07, GJRR08, GJRR11, GJRS12, HJRS99, JKRR04, SSJR10, CJRT05, DJRZ06, DJNR09, GJRT01, HJRS99, LRS98, Rol97, RV02]. Randomized [Die04, Tei93]. Randomness [Gol11b, vM00]. Range [Yu02]. Rannoch [CM00]. rapid [GH93, Gue05, GS06, GB07]. Rapperswil [DGJ09]. RASTA [LMP04]. Raster [HAB98]. Rationale [Bar08b, DHM99, Bar97, CW02]. Rationality [HHP09, vDLJ11, Bv97]. Raton [BTMI11, BHLP07]. Ravenna [MF11a, MF11b]. Ray [dB93]. RC [DW12]. RCG [BEM06, Lag05, MH05]. RDDS [MTH06a, MTH06b, MTH07a, MTH07b, MTH08]. re [Lou94]. re-engineering [Lou94]. Reachability [BP09b, DP11, FLP12a, KP10c]. reaction [CM95]. Reactive [Bee00, BJK^{+}05, AW05, BR97, LL95]. Reactivity [GHI^{+}06, HWP01]. Reader [Ceb08, DB98b, Eug09, HM02, SC07, BM98, BBC04, Fro02, Kow01, MMM01, MØ05a, MD99]. Reading [AvAS06, AvASD06a, AvASD06b, AvASD06c, BHT97, DJW04]. Readings [vdV84, Ram98]. Real [AH98, BR97, BCDN07, CLS07b, DO02, DFP^{+}12, FT04b, GKL^{+}11, HHLR08, JLW07a, Jos00, LY04, Rei99, AW05, BC04a, HWP01, Hoo91, JP96, Jos88, Kat99, LdRV94, Mal97a, RR98a, Ric02, Vyt91, Cag00, CH04, KPR07, MVCM10, dB92]. Real-Time [AH98, BR97, BCDN07, DO02, GKL^{+}11, JLW07a, Jos00, LY04, AW05, BC04a, Hoo91, JP96, Jos88, Kat99, LdRV94, Mal97a, RR98a, Vyt91, CH04, KPR07, dB92]. Real-World [DFP^{+}12, Rei99, HWP01, MVCM10, Cag00]. Reality [BST03, CD07, CHCM11, DSL08, HKP10, LEP^{+}10, PCH^{+}06, Shu07, Shu09, Shu11a, Shu11b, Sub05, TGM97, YJ04, YJS^{+}06, Aya95, BST01, BKS96, ÁCD^{+}12]. Reality/Augmented [CHCM11]. realization [PD96]. REALWSN [MVCM10]. reason [Kra97]. Reasoning [AW12, ABMS06, ABMH08, ABD08, AL11b, BV02, BS11b, BGLT05, BBS07, BKLM07, BR04, BGP08, BGP11, BM10, BV12, BM11b, BHM03, CBC^{+}02, CL08a, CVV08, CV10, CP02, CRW^{+}12, DV09, DF11, DV07b, EF02, EH08, ELS09, ELLP12, FV10, FNGW08, FS06, GW09, GH10, GMS12, Har11b, HV06, Heu04, HL10, KI01, KS12d, Lam00, LSR03, LBSBW98, LN03a, Liu11, MPdSM07, MP03b, MW09, Mel07, MAR05b, Nar01a, NV01b, OS04, Oli07, PWZ98, PV00, PS09b, RW11, Ren02, RBGG06, RG11, Sch98b, SKP07, SC09b, VV03, WR07, dCdF^{+}08, AW01, ABB99, AGT98, AB03, BV05a, Bau98, Bec05, BB01a, Ber99, BP01, Cad95, CKM93, DFN97, Dyc00, EFT01, FS05, FCF92, FKKB^{+}05, FK95b, FG04b, GO96, Gab97, Gal97b, GVM99, GLP99, God05b, GLN01]. reasoning [GH92, HKM95, HP99c, HS05b, Lau93a, Lau93b, LP97, LCB97, MNT95, MP94, Mur99, NZ03, Pfe94, Rei89, SF96, SC98, SV05, VA95, Vor92a, Vor93, Wag94, Wat95, WAR94b, de 98, AP96, AAB^{+}07, ABW10, BBF^{+}06, BBM^{+}08, DvV96, EM05, EK12, Neb90, PdA^{+}11, Sch03b, SS06e, TFE^{+}09, LN94, Vel94a]. Rechensystemen [Lei74]. Rechner [BHH74]. Rechner-Gestützter [BHH74]. Rechnerstrukturen [HS74]. Recife [BR02, BCSW10, CSW06]. Recognising [QCDM06]. Recognition [AMGJ12, AMPT09, BCE09, BC10b, BK02, BKLS11, CAD^{+}02, CK04a, CK04b, CK06b, CK06a, CK08, CK10a, CK10b, CK12b, CK12c, COMTLB12, CMM^{+}09a, CNA08, CD00b, CBBS09, DNS09, DTMH10, DJ08, EV07, FMNS06, FCD^{+}04, GDP07, GHI^{+}12, GRK^{+}10, HSJ07, HWW^{+}10, IS12c, JFT11, KSG^{+}09, KC07, KC09, KC11a, KC11b, KMKP11, LV02, LLT^{+}04, LST^{+}05, LL06, LLO08, LK04, dVLKR^{+}08, LWRdR11, MST12, MdlBP05a, MdlBP05b, MBMS07a, MBMS07b, MTOK06, MTCOK10, MTCOBH11, MLTC11, MF11c, MK03a, OLL10, PH11b, PCPS03, PI05, Per07b, Per09b, Per11b, Per12b, PPBL12, PMS08, RWA06, RSV07, RFZ03, RBGS04, Rig08, RMK07, RSK08, SK11, SRS03, SM06, SE10, SK12, SKSA12, SSAP05a, SSAP05b, SLCT11, VSH11, YKF^{+}06, YZCW07, vG02, Ami98, BV05b, FIAP01, FZJ01, HP99b, HV03]. recognition [KC05, KT96, Kit88, KSH05, PBB05, PH97, PpWR96, PP99, Per01, PR03b, RF01, RVY05, Ruc96, STC04, SL05, SMK01, TC98, TEB09, DJW04, PHSZ06, Can89, Gor02, Tan93]. Recognizing [Rei93, ÜÇA10]. RECOMB [AGI^{+}06, BS11a, Bat09, Ber10, BEM06, Cho12, CM09b, EW05, EIRW06, IB07, Lag05, MH05, MMK^{+}05, NV08, SH07, Tan11, TD07, VW08, IWC04]. RECOMB-CG [CM09b, NV08, Tan11, TD07]. Recommending [DLN10]. Reconfigurable [HG00, GZR02, BWAM09, BCV06, CCAS12, DMB^{+}07, KKM^{+}11, SMEGA10, WCBD08]. Reconfiguration [Han02]. Reconstruction [Heu04]. Recover [Pfl02]. recovery [Don89]. RED [Lac10, LV12]. Rediscovering [Cai11]. Redmond [CJ11]. Redondo [BA07, Jen01, KP04]. reduction [FK87]. reductions [BKR87]. Redwood [KKP02]. Reference [Mos04, TDB^{+}06, Uni81, Bee86, Lis81, Luc87, TD97, TDBP01, UA83]. refinement [MO97, dBdRR90, CSW06]. refinement-algebraic [MO97]. reflection [Coi99, YM01, CST00]. Reflection’99 [Coi99]. reflections [dBdRR94, BCK08]. RefMod [tHBP08]. Reformulation [KH02, MR07b, CW00, ZS05b]. REFSQ [BF11, GH09, PR08, RD12a, SPH07, WP10]. Regensburg [EBT07, LPT07, MVC07, PW07, SEN07, WRP07, WSG07]. Reggio [SC09a]. Regina [SYP^{+}05]. registers [Ron84]. Registration [DCFR12, FDL10, GMV03, PLG06]. Régua [CBF09]. Regularities [Pup10]. Regulated [BPD06, LMP04]. Regulatory [EIRW06, EW05]. Reims [MCdS^{+}08]. Reinforcement [GLM^{+}08, SH12a]. Reinhard [RSB07]. Reinhardsbrunn [AJ94, BJS93, Jan89]. Reis [MEB05, RM04a, SW07]. Rekursive [BHR84]. Related [BM11b, EF02, GW09, MP03b, OS07, Oli07, AP93, Ant97, BHP95, Bec05, CKPB07, Cha03a, Chu03, Dyc00, Gal97b, Kim01b, KMTU03, MMMO96, Mur99, Sch92, SILS05, de 98]. relation [Lev92]. Relational [BMS04a, BMS08b, BJM09, Jae01, MWD06, Sch06, dSOSR04, dSOSR06, BC89, Fre87, Gra91, San80a, BMS04a, KG12, MWD06, Vos92, dS02, dS11a]. Relations [Geu98, AFS89, Bol99, BMS08b, BJM09, Sch06]. Relationship [EKT94, Lou94, PT92, Dau03, Gog94, Pap95, EKT94, Lou94, PT92, Zwi89]. Relationships [SG07, BRR87a]. Relevance [PNB04, SNZ08, NSM00, SVV12]. Reliability [AFL03, BRS09, FBV11, Gór06, HS08a, HLW04, KvdM00, OD12a, OD12b, QBY^{+}12, SO07, Sch10b, TQYX11, ABM02, Bit88, Ehr98, FKP99, Vog01, WGD05, ON80]. Reliable [AK07, Asp98, BP12b, BKM08b, CS01b, GdlP99, HB97, KV08, LS04b, PH06, RV10, RV11, RS03, Str96, VW05, BS02b, AK07, Asp98, BS02b, BP12b, CS01b, GdlP99, HB97, HHLR08, KS07b, KV08, KK09b, LS04b, PH06, PK00, RV10, RV11, RS03, Str96, VW05]. RelMiCS [BMS08b, BJM09, Sch06, dS02]. RelMiCS/AKA [BMS08b, Sch06]. Reloaded [dMGFS11]. Remo [BDV95]. removal [dB93]. rendering [The89]. Rennes [BFT^{+}08, BFK^{+}08, KBP07, Kot86, KB92, eOP05, Rao92, VA07]. Repbulic [BTQ03]. replacement [Hab92]. Replication [CBPS10]. report [AS02b, BP95, BJ95, HR91, JW74, JW75, KSW90, MO92a, PBB^{+}83, RMV96, DMV01, Smi03, TSE^{+}09, BFS92, EDD04, JL95, MPS93]. Reports [Cha09, DS11b, Gho10, Gie08, Kie12, Küh07, Ceb08, Eug09, SC07, ÜÇA10, BBC04, MØ05a]. Repositories [PGG^{+}06]. Representation [BS11b, BPCS06, Bar02, BGP08, BMS04b, CPR12, CRW^{+}12, GMS03, GJN10, GJTS06, LSR03, MB09, RtTMP10, RtTMP11, RtTM12, SHL08, AGL06, DPP98, EEM96, Fre03, Heb95, HNM02, Her94, LN94, MP94, MMS93, Mor89b, Neb90, PZH96, PB92, Ste12, dPS95, GTWC09, Haj96]. Representations [EV07, JFT11, AGT98, BV05b, HV03, Tau89, TEB09]. Represented [BBH^{+}10]. Representing [FHW98, Had94]. Republic [BM05b, BSW95b, BE87, BBK12, BKL^{+}12a, BKL^{+}12b, BFG^{+}12, BTPP^{+}12, BVNW87, BJKK02, BJ12b, BR07c, CUDL02, CCMR10, CKMS05, DV10, De 07a, DMWW88, EV09, FKK04, FKM09, FPCK11, FN12b, Gie10, GSS09, GLM93, GV95, GGS99, GP02, GRP10, GDGB12, HM11a, HKR07, HS94, HKP10, HS95c, HJW00, HMP99, HK10b, HŽ07, HSH08, HS98c, HL91, HdL12, HKM^{+}11, Jay99, JKB96, JJP^{+}02, KMW03, KS01, KPP08, KK11d, KBČ^{+}12, Kra99, KLFK10, KNK08a, KNK08b, Kuc̆02, KK07, KP10c, LD97, LMŠT01, Mac06, MP03a, MJ04, MRŠ03, MMP03, MMV04, MMOS99, Mat01, MM03, MMP05, MM07, MM09b, Nes99, Nic86, NKM^{+}09, OPT01, Ott87, PR01, PM04a, PM04b, PM04c, PM04d, PTRH05, PP12a, PTB99, PJ97, PBRB08, QBY^{+}12, RRBE09, RW86, SLH04, SPK01, SKP00, SKP02]. Republic [SKP04, SKP06, SHKP08, SHKP10, SHKP12, SKK04, SS06c, Ste99, ŠPTM00, TS87, Tra04a, VPBŠ03, WH95, WTP^{+}06, ŽL08, ZBG98, ŻR99, vWN99, vLIvdH^{+}07, vLMP^{+}10, vW97]. reputation [FBKS03]. Requirement [PM08]. requirements [DEM81, Yao82, BF11, BR98, GH09, PR08, RD12a, SPH07, WP10]. Reseach [UMN^{+}08]. Research [AW12, ACR07, ADMH04, ADM^{+}06, ADM^{+}08, AGS^{+}11b, AGS^{+}11a, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, AD11, AJ01b, BN09a, BYdVZ^{+}12, BCT02a, BMUV10, BMU^{+}09, BHHK08, BEJvR11, Ber96, BJSW09, BM10, BG12b, BL07a, BMSW12, BGPR10, BF10, BBMSD09, BL04b, BDFS04, CKD11, Cav85, CZ05, CWZ11b, CCNY10, CGG^{+}07, CFG^{+}11, CKV08, CGvR02, Cup00, DEQ92, DFPS06, EDD04, ECB06, FYM12, FKM07, FT04b, GLZ^{+}03, GCL^{+}06, GT03, Gol11a, Gol94, GKW02, GMS06, GPE05, GNTF02, GS08, GPT10, GHK^{+}10, GDG11, HW07, JSV11, JL08, JDAW12, KGL04, KEM07, KMS06, LK09a, LMR^{+}06, LK99, LGR^{+}04, MOP^{+}08, MP06, MSG07, MG08, MZ07b, MSZ08, MNZ09, MM01, MSWZ06, MT04b, MW09, MP93, MAR05b, Nie94, NS81, PPPM10, PRK12, Pro11, QBY^{+}12, Qui98, RW11, RLL04, Sch03a]. Research [SF08a, Seb03, SCP^{+}12, SB12a, SCH04c, SG03, SD06, TC95, TBC07, WK08, WR07, WZA10, YAKM^{+}12, ZRP^{+}09, ZRP^{+}11a, ZRP^{+}11b, ZYBX08, ZYW08, Ada96, AY96, AW01, ABB99, AB03, ABGP05, BP05, BS05, CE95, CGLR05, DLW06, EJS98, GLIC10, HRR93, KJ91, LP97, LFL05, QS00, Ric02, SRG04, VA95, ZTY^{+}05, Zho01, ZLS^{+}06, dSG05, AV99, AT02a, ABK^{+}09, AGI^{+}06, AL12, BS11a, BHH^{+}04a, BSC12, BL92, Bat09, Ber10, BB00a, Cho12, CDCJL08, CM00, CS01a, GTVC06, GTW02, GBMS11, HL04, HKP10, KS04b, KFM07, LJR^{+}10, LSW08, MMK^{+}05, NS98, PT97, RCT05, SS83, SH07, VW08, WKK98]. Reservation [SF10]. Resilience [ADB^{+}12a, ADB^{+}12b]. Resilient [Avg12, Tro11, HS08b]. Resolution [Bra07, DBM09, HDBM11, Bur91, Fra96, San80b, Sch97b, Fer93]. Resource [CC12, PvES12, SF10, vES10, Lac10, LV12, SWB06]. Resources [DLdS04, TO08, SF10]. Response [LK07, LK09b, Bee86]. restricted [Bur91]. Result [AFKL04]. Results [BMU^{+}09, HS12a, KMS06, LK09a, NM08, Rei99, SG10, WH11, Mau91, Sch97a, SKMR94]. Rethymnon [BBP^{+}12a, BBP^{+}12b]. Retirement [KR11]. Retrieval [ACR07, AC11, ATM09, AKR^{+}09, BYdVZ^{+}12, BHL^{+}03, BBMSD09, BDN08, CBGCCZ12, CCS04, CMSM^{+}10, CL10, CKLN10, CFG^{+}11, CBS02, CFS06, CHR10, DJNR06, DLN10, DGSN11, DKN^{+}11, EKO^{+}04, FLMK06, FLT07, FKLT08, GKT09, GKT10, GKST11, GKS12, GSS11, GHK^{+}10, HRdV11, HNS^{+}12, KTH09, KMVY06, KMYJ08, LO02, LMR^{+}06, LSL^{+}09, LSE02, LLM^{+}08, MOP^{+}08, MMBND07, MT04b, MGSM12, NdMO03, NLKJ06, ND04, PBGK03, PCG^{+}07, PJM^{+}08, PDK^{+}10, Qui04, RBSL06, SL12, SSO^{+}11, SNSR06, Wii04, YKMJ09, ZBY07, ACP01, AM04, BCBG05, CN05, CGvR02, FLMS05, IS98, LYMM05, LLC^{+}05, LFL05, MZWZ05, Pet01, PBGK02, SS83, Seb03, Wii05a, Aso04]. Reunion [KR02b, PV00, LD05b, LD05a]. reusability [Fra00]. reusable [US90]. Reuse [BK04a, EK09, Mei08, Mor06, Sch11, Fra00, Gac02, Fen00, Mor06]. Reverse [Che99a]. Reversible [DW12]. Review [LL06, BLW87, SG96]. Revised [ARS03, AJS05, AKK04, AS11a, ABG^{+}06, ABH12, AGP08, AM03, AdBB12, ABF^{+}07, AKU01, AK03, ABK05, AKT06, ACK^{+}07, ABKT11, AMU03, ADB^{+}12a, ADB^{+}12b, ABMS06, AV07, ACW07, ABdJS10, AKR^{+}12, Alp11, AFKL04, AM07, AC09, Ama08b, APW06, ABJ02, AT07, APN08, AMNT06, AS10, AÖ12, AS12, ABGL09, AK12, AS07b, Ari02, AL11b, AL12, AOSV08, APV09, AAR05, AJ01a, AHSC11, ACHH12, AJN12, AKS09, ABRS06, ABFR07, BBMB^{+}08, BBD^{+}10, BGHJ07, BW03a, Bai05, BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, BS09, BJ10, BFGM05, BYLJ11, BL12, BKLM07, BK05a, BHKR11, BHJKPG06, Bar03, BBH^{+}05, BGH06a, BF08, BDE12, BRS^{+}07, BPR05, BBP^{+}12a, BBP^{+}12b, BM11a, BDG12a, BBDS12, BHGP02, BS06b, BDG^{+}12b, BB05a, BCD04, BDd09, BHM^{+}04, BHMS07, BCV06, Ber03, BLSS11, BVAT12]. Revised [BY07, BZU07, BGS11, BBZ01, BS10, Bla03b, BJN03, BPT06, BRST08, BC05, BDC12, Bod03, BRU06a, BHU03, BFSN10, BPD06, BCC^{+}12, BMN05, BKW09b, BJ01a, BLvdH10, BB06, BFMS08, BFJM09, BCSW10, BDDS05, BDDS06, BDDS07, BC08, BG07a, BGL09, BGJ11, BR07a, BFT^{+}08, BDN08, BRM06, BvLR04, BKL^{+}12c, BKM01, BGKS04, BBPP09, BDGN07, BR07b, BC11c, BKM07, BBT10, BB11, BGHS05, BBK^{+}11, Bre03, BAdVR08, BHLvdP07, BEFP08, BZ03, BKM06, BKM08b, BtHRBB12, BSH06, BHJY07, Bru06b, BW09, BS12, BS01c, BBB06, BKB03, BR07c, BY06, BTMI11, BDLMM12, BHM^{+}02, BFOY04, BTF05, BH06d, BCDN07, BJ01d, BGH06b, But10, BGW06, CJ11, CG12, CCS04, CG04, CKP^{+}12, CCJS07, CKD11, CL11, CK12a, CRL01, CV04a, Can12, CGL^{+}09, CBG^{+}10, CBS^{+}12, CHB^{+}12, CV08, CMSM^{+}10, CP12]. Revised [CS03, CS05a, CPSA10, CGMP11, CSW06, CA08, CAS^{+}09b, CW11, CY03, CMZ99, CM03a, Cha05b, CUDL02, Cha03a, Cha09, CWiM^{+}09, CR04a, CH04, CPZ07, CLZ^{+}09, CF09a, CY10, CY11a, CYZ12, CHG^{+}07, CM09a, CHZ08, CBS^{+}11, CEFZM05, CKO01, CK06c, CS10, CSV05, Chr01, CCMR02, CCMR03, CCMR05, CCMR06, CCMR07, CCMR09, CCMR10, CMMR11, CCMS11, CMSA12, CSN14, CCG08, CSY09, CY11b, CW01, CGP03, CFVV^{+}12, CFVCF^{+}12, CV10, CVV12, CPvdH03, CF09b, Col07, CML^{+}10, CDN12, CR01b, CLS07a, CM00, CMCS11, CFP^{+}09, CM09c, Cor02, CKTW12, CF10, CS12c, CvRVSN12, CMOZM11, CDM04, CRW^{+}12, CGBDA08, CLBP11, CKLM09, CE00, DFL^{+}10, DR02a, DKY06, DYCD09, DGT10c, DM06, DG06a, DDS12, Dan12, DF10, DDE04, DSLT08, DSRW08, DSLT10, DSHL11]. Revised [DS09a, DM01b, DLT05, DSS11, DDHP05, DPC^{+}07, DF09, De 07a, De 10a, De 10b, DFPV11, DW12, DS09b, DHtM^{+}12, DGG^{+}11a, DGM09, DV09, DG10, DEG11, DFP^{+}12, DBB12a, DLW12, Des09, DJS06, DJNR06, DLN10, DGSN11, DKN^{+}11, DR06, DR09, DP10a, DGWB09, DPA07, Die02, DD07, Die03, Dig04, DvEF06, DGR06, DGDM^{+}11, DMRS06, DMRS07, Dim03, DDK11, DS11b, DS03, Din03, DG09b, DKN12, Dit01, DFN10, DB07, Dol09, DO11, DS11c, DFT09, DMW06, DW06, DW07b, DKR02, DKN02, Dun09, DS06, Dur12, DS07, EJM10, EDGP11, EL07, ELS12, EKF12, EKO08, EM05, EKP^{+}07, ET01b, EPVW10, EG10, EP06, Erl06, EK07, ENO12, Esc10, EIRW06, EFZKM07, EBAH08, EV09, EHMM09, ECV^{+}10, EVV^{+}11, EEM^{+}11, EH12, FRS08, FBSMS08, FV01, FV03]. Revised [FPFR05, FPRAW04, FS07a, FZJE^{+}05, FZJE^{+}06, Fei03, FR01a, FRS02, FRS03, Fek08, FKK09, FGB11, FLVS11, FMZ05, FL09, FMO05, FS07c, FKM09, FdCLJ^{+}11, FdCLN^{+}11, FPCK11, FST09, Fom06, FLUV02, FS07d, FS08, FS09b, FS10, FL11, FKM08, FN12a, FKKB^{+}05, FPRS06, Fuc10, FLMS05, FLMK06, FLT07, FCDB07, FMY04, GVO^{+}10, GKL09, GML^{+}08a, GPCL10, GANACBR10, GANACL11, GANACBdCdV12, GAL12, GCFA12, GVCA07, GY08, GLV09, GMV03, GKS^{+}07, GMP06, GKT09, GKT10, GKST11, GKS12, GG02a, GHP^{+}11, GPR^{+}12, GGS09, Gho10, Gib12, GCK06, Gie08, GKL^{+}11, GH05, GH12a, GH06, GMO03, GHSW04, GLM^{+}08, GT04, GOZ05, GGS11, GSS09, GS05, GGHS07, GSLM12, GK02, GR06b, GN05, GP08b, GCC^{+}02a, GHMT05, GT05, GW06b, GNTF02, GF12b, GO06, GKMN10, GHI^{+}06, GM05, GVT^{+}11]. Revised [GAR03, GAR04, Gue05, GRR05, GS06, GB07, GM07b, GDG11, HM11b, Hag02, HO03, HENR03, HS12b, HH04, Han09, HLC^{+}12, HK12, HKI^{+}09, HNRW06, HSSY05, HDBM11, HM03, HBN06, HPR05, HHLR08, Hic04, HBH08, Hig05, Hil06, HRT^{+}03, HRR^{+}06, HPS09, HR08, HCF06, HhKRT06, HCR06, HŽ07, HNQ08, HI10, HPRS06, HM05a, HMMW10, HNP12, Hor06, HZB07, HPSZ08, HPZ10, HNW04, HM11c, HNPP07, HGBF12, HGGG12, HGPH12, HMBD09, HKS09, HHG11, HDLM09, HMSU04, HdMKS12, HKM^{+}11, IB07, IPSK04, IS11a, IS11b, IST07, IJ06, IGN05, IHX^{+}08, IDMY08a, IDMY08b, IWC04, IT03, IKKU08, IS12c, JRSN09, JMBS07, JSS08, JSO04, JSO11, JP07a, JLST06, JNS05b, Joh03, Joh01, JP11a, Jón12a, Jón12b, JS02, JJF03, JDMB08, JDMB10, Jou11, Jue04, JP12b, JY12, JRR03, KEDW07, KS03a]. Revised [KG04, KS08a, KN09a, KKP02, KCR04, KKS11a, KKS11b, KA12, KUA00, Kap08, KS09b, KW07a, KM08b, Ker12, KEM07, KS12b, KL12, Kie12, Kim05b, KYL07, KSB11, Kim12, KKS11c, KKS11d, KKS11e, KKS11f, KM04, Kin08, KDS01, KHG01, KH11a, KH11b, KMSS02, KTB09, KK11d, KBHSW06, KHSM^{+}08, KS05a, KRMM^{+}02, KPS09, KW10, KKC09a, KKC09b, KS07a, KS07b, KBČ^{+}12, KMS12, KS05b, KLT^{+}07, Kow01, KMTU03, KH11c, KMVY06, KMYJ08, Kuc̆02, KBC^{+}09, Küh07, KRS02, KZ10a, Kur10, KCO04, KŽ10b, Kut12a, Kut12b, KCK08, KBC05, LSJ06, LJS08, Lac10, LV12, LRS03, Lag05, LYL11, LSV06, LVS08, LS10a, LSW01, LSW02, LSW03, LO11, LN04a, Lau12, LS04a, LT11a, LL03, LhK03, LYY07, LC09, LH10, LY12, LMSS07, LOST04, LT05, LOTY05, LS05, LM08a, LWM^{+}08, LLLL08, LLP^{+}07]. Revised [LSDN04a, LSDN04b, LOS05, LVW05, LLM^{+}08, LCF^{+}04, LL04, LY05, LAF^{+}10, LMP04, LMCH12, LYCV08, Lio04, LMWY04, LMW08, LMW10, LMW12, LN01b, LCY03, LA05, LLL06, LL06, LW12, LK04, LK07, LK09b, LOH05, Lop06, LH08, LS06b, LSU07, LP08, LGS09, LSW08, LS09, LV07b, LCY^{+}11, MWD06, Mac12, MMMP12, MP06, MSG07, MG08, MRK05, MBS11, MY06, MMBND07, MS06a, MWY01, MVW09, Mar05a, MŠKL02, MOBS06, MT02a, MSY08, Mar11c, MRU12, MF01, MVMP^{+}04, MVOF08, MS05, MP10b, MR12a, MS00a, MTV09, MPT10, Mat02, MZ04, MS03b, MPR07, MPPJ^{+}05, MRY^{+}11, MRPM10, MRP11, MPR12, MMM10, MFG12, MLTC11, MCSN10, MT02b, MJ07, MNW08, MW08, MB09, MN12a, MN12b, MSH08, MMG^{+}01, MCSY07, MV12a, MA05c, MNLM06, MP12, MOT12, MTC^{+}08, MSB07, MvL10, MS11a, MF06, MG07b]. Revised [MDBF01, MDPQA09, MDPQA12a, MDPQA12b, MBS05b, MK03b, MBJ^{+}07, MK12, MS02c, MOTN07, MTNL12, MPS04, MZ06a, MGSM12, MNTY05, MSS05b, MJL02, MZWZ05, NMM10, NP09, NP11a, NP12a, NZZ11, Nar01b, NZS^{+}06, NSS^{+}07, NOOS05, Ngu06, Nie09, NR06, NLS08, NVSB^{+}07, NG10, NvRSS11, Nür03, ND04, NH03, Nyb08, OROD07, OO12, ONPR07, OFR09, OGM05, Oka03, OS07, OBS12, OW09, Ölv10, OTZ01, OPT01, OPP04, OSV11, OBM11, OSRS07, Pac05, PSP^{+}02, PAV^{+}10, PZ07a, PM08, PHZ11, PP01, PVW10, PTRH05, PR07, PT09, PJC^{+}12, PAD^{+}08a, PDML11, PPCS12, PH04, PC05, PMMR06, PSPH^{+}10a, PSPH^{+}10b, PS12a, PH10, Pau03, PPJRN^{+}10, PYLW08, PW09, PA03a, PvES12, PP12b, PRT06, PSO05, Pet01, PBGK02, PBGK03, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, PK03]. Revised [Pet03, PNPP12, PU04, PTZ03, PNB04, PKSS02, PVV10, Pol01, PBRB08, PR98a, PT06a, Pri03b, PQ05, Pro11, Pue07, PS09c, QLXL07, QCDM06, RMR05, RPR08, RM09, RLL04, RUL04, RRU05, RTZ08, Ras01, Rau04a, RKM^{+}12, RTB02, RG03, RBF06, RSS06, RN11, Ria08, Ria09, RtTMP10, RtTMP11, RtTM12, RK09, RGW01, RBTD04, RL11a, RSL01, Rob06, RH01, RHD08, RB10, RF12, RBSL06, RFBD09, RM04b, RME04, Sad13, SS08a, SNY06, ŞS05a, ŞSW07, SM11a, SSVW12, SFS11, SHN07, SÇS12, San02, SH12a, SINK08, SGG06, Sav06, SMB03, Sch10a, ST08, ST11, SKP07, SKP10, SC11, SJDT08, SNRG11, SSN08, SNZ08, SVV12, ST12b, SHGB05, SHW09, SG09, SDH05, SCL^{+}06, SLL^{+}07, SYY^{+}08, SPÖ^{+}10, She03, SXC04, SKPG12, SBD03, SA06, SPON08, Sil01, SM11b, Sio10]. Revised [SCD^{+}10, SA12, SHP^{+}05, Smi05, Smi06, ST07, SSM08, SCZ10, SOP12, SMF^{+}11, SKK04, Squ09, SKCK07, SP07, STWS07, SG07, SBF08, SZ11, SM09, SHŚ^{+}07, SSZ12, TvdBI02, TT10, TBC07, TNA^{+}10, TKSN11, TPCB12, Thi10, TGH06, TP09, TT06b, TST09, THR03, Tsu08, TYNM05, TtHVS06, TWSN07, TNGK08, UWB^{+}06, UBW06, VS09, Van05, VAR12, VMvN12, VEG^{+}11, VY01, VFC08, VU09, Vet11, VHM07, Vid12, VSMS12, VV07, VS05, VASP09, VKG12, VWY01, WS02, WK08, WR01, WmCL07, WJL^{+}12, WZH^{+}07, WSTI07, WW03, WPM05, WPM06, WPM07, WBD08, WMdLA10, WG11, Wii04, Wii05b, WL03, Win04, WNL05, WKB04, WHR10, WPZA11, WK06, WY98, WVDS05, WCGB05, WKD08, WDPW02, WDPW04, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, XW11, Yal10, YLWY06, YJKK06]. Revised [YJKSW10, YC11, YSL12, YMY^{+}10, YY09, YKMJ09, YAKMJ10, YAKMJ11, YAKM^{+}12, Ytr06, YP01, YLL09, ZSD03, ZSSM03, ZiTM10a, ZiTM10b, ZiTM10c, ZSM^{+}09, ZCDT10, ZZZL12, ZLY^{+}07a, ZY01, ZHP12, dCdF^{+}08, dBBGdR03, dBBGdR06, dBBGdR07, dBBGdR08, dBBM09, dBBHL10, dGEK11, dGN12, dKIM^{+}07, dMGFS11, dS02, dSOSR04, dSOSR06, tCZ07, tHBP08, vDKS11, vES10, vEHD05, vEDA04, vKS12, vLS11, vdBK05, vdBGG10, vdHHsHD06, vdHCD07, vdHS10, vdHIP11, vdHP12, vdL02, vdL04, vdMS11, AMW07, ACP01, A^{+}08, ADB^{+}05, AKR03, AGL06, BCG02, BDGM05, BR06, BMS04a, BC04a, BK05b, Bir07, BL06c, BHK01, Bru04, BD03, BT06, BS06d, CP06a, CHPW05, CR05a, aCF06, CZ05, Cho02a, Cho02b, CKMS05, CMT01, CUdlC04, CKL05, CRYS09, DS05a, DS05b, DPA05, Dik04, DL04b, DRS05, DOSY05, DLW06, ELM05, ÉEMS04]. revised [EW05, Eta05, FV04, FV05, FRA05a, FLS06, FRS05, FFRS05, FPMW06, Fri04, GOW03, GAN05, Har01, HN06, HRTR05, HW06b, IMW04, JSdSA05, JJ03, JY14, Kah03, KIRK05, Koh06a, Kra05, LBCO04, Leu03, LBO05, LMT^{+}04, LMW06, Mac06, MML97, MNS05, MEK03, MS03c, MTW05, MDP04, MSS05a, MBA05, OFK05, PY05a, Per05, PGBK04, Pie05, PT06b, RWY97, SLB05, SKY06, SS04, SS06d, STLK05, SHO99, TLC06, TBG05, TMP04, TAS05, Vää99, VZ06, VU05, Wii05a, WPH03, WWC02, Wri03, dBBG04, dBBGdR05, dRLP98]. Revision [KI01, Neb90]. Revolution [BDFM12]. REWERSE [BM09b]. Rewriting [Baa07, CDE^{+}07, DHMM96, Gie05, Pfe06, CJ95, DL95a, Dor95, Hee94, KJ88, KO91, RR93, Bac00, Boo91, CLKK07, Com97, Der89, Dur12, Gan96, Hsi95, Jou85b, Kir93, Les87, Mid01, NR99, Nie03, Nip98, Ölv10, Tis02, Tre09, Vor08, vO04]. REX [dBdRR90, dBdRR91, dB92, dBdRR93, dBdRR94]. Rey [GYA^{+}06, Ish97, KSHM09, PISW05]. Reykjavik [ADG^{+}08a, ADG^{+}08b, BKR09, BFGL10, BD11, DR11, EH12, Fek08, FR11, KL96, LRH10, RBB^{+}06, VKMT11, Jón12a, Jón12b]. RFID. [JP12b]. RFIDSec [JP12b, Yal10]. RGU [BHH74, BHH74]. Rheinfels [BHP95]. Rhode [SRW98]. Rhodes [AMSM^{+}09a, AMSM^{+}09b, Dol09, KPS12, Wim03]. RI [Bes07, HKS09, Ish11, LP97, DST01, Sil01]. Ribeirão [SCR06]. Rica [dVGR04]. Richmond [Li11b]. Richter [Len04]. Rico [CMM93, Rol99, RF99b]. RIDS [GB97, Sel95]. RIDT [HAB98]. Rietz [BJ86a]. Riga [AF06, BDGM05, BGHS05, Fre01, GMV09b, GKMN10, MA07c, PS04b]. Right [EL02a]. Rights [BBGR03, SNY06, Fei03, San02]. RIGiM [CVLL12, HRK^{+}07, HP09, SPC^{+}08]. Rigorous [BBPP09, AG09, BJRT06, WH11]. Rijn [HM11b]. RIM [ANS04a, ANS04b, ANS04c, AIL12, BK05a, BHH08, CL06a, CCH02, FG96, GGS09, HK05c, HZ08, Ho12a, Ho12b, HhKRT06, HXC^{+}08, IAL^{+}07, Ish99b, IS02, KL02, LM98b, LB03, LXH^{+}12, LS09, MR07a, MS00b, MWK^{+}09, NcC99, QLK^{+}10, QLK^{+}11, RK09, RK12, SS06b, SLC01, WHL09, YW06, YY01, ZS00, ZGY04, ZO10, ZYRH06]. RIMS [Got83, GAY86]. Rio [BPAB07a, BPAB07b, BC09b, CDGW10, CEM^{+}06, DPC^{+}07, ERK07, ES03, FMS10, LdQ07, LSHB07, SMK01, ZZTG07]. RIPE [BP95]. RISC [Knu99b]. RISE [Gue05, GS06, GB07]. Risks [NRW04]. RISSEF [WKB04]. Riva [EL07, FdS12, NVSB^{+}07]. Riviera [Gil10]. Rivières [ELN^{+}10]. RLCPS [DDS12, SCD^{+}10]. road [MvOvRdV05]. Roadmap [HG00, BS05]. Robert [KS02a, KS02b]. Robo [VPK00]. RoboCup [AK99, BLNG10, BCT02c, BJNT06, HWP01, IMWZ09, KLR03, Kit98, LSST07, NRSSV05, PBBY04, RMSS12, dSCP11, SBK01, VPK00, VROD08]. RoboCup-97 [Kit98]. RoboCup-98 [AK99]. RoboCup-99 [VPK00]. Robot [AK99, BLNG10, BCT02c, BJNT06, IMWZ09, KLR03, LSST07, NRSSV05, Pau01, PBBY04, RMSS12, RHD08, dSCP11, SBK01, VROD08, Kit98, WD00, KPS01, SK08]. Robotic [Bee02, GAM^{+}11, Har11b, BHGP02]. Robotics [GLCT10, GKC^{+}12, HSP^{+}12, JLS11a, JLS11b, KGV^{+}09, LDXZ10a, LDXZ10b, MBH^{+}11, ŞSW07, SRL12b, TGM97, WPE05, XXX^{+}09, XHXL08, XLHX08, Aya95, BL89, DvV96, DTF98, Gom01, HM98, Mit98, ŞS05a, SRL12a, SRL12c, Don89]. Robots [ABH^{+}10, CNP^{+}08, NABK12, WK08, WPE05, WD00, BD98, CBN99, Hag02, Bee02, Har11b]. RoboWorld [HSP^{+}12, KGV^{+}09]. Robust [FMS02, Nag01, AMZ09, Gor02]. RobVis [KPS01, SK08]. ROC [CH04, Wan04]. Roche [FS07c, Gue90]. Rochelle [OLL10]. Rochester [Dwa00]. Rockland [GS94]. Rocquencourt [APFV09, FRS08, GP07, GP09a, GP11, HKS09, KG74]. Rodney [DDS12]. Roever [DHS10]. ROIA [CAS^{+}09b, LAF^{+}10]. Rolduc [Nag90]. Role [AN06, CKH10, CR05b]. Romania [CP99, Gel10, GNTF02, HPSZ08, Pau03, PPJRN^{+}10, PLK^{+}12]. Rome [AT02a, ACR07, Aus86, BP07, BW03b, Ber96, Boh75, BBD97, BGP00, BCP94, CFI06, CD10, CGS01, CFH^{+}12, CMMR12, Coe11, CGS02, CR73a, CR73b, CP05, Dem07, Di 97, DSV01, DFF06, Ehr04, FJ10, FS09b, Gia89, GG02b, JLV02, Jon01, Knu99a, Len97, Leo04, Leo09, MSS11, MB01, MP03b, MR02, Mor89a, NBC^{+}06, OFR09, Pel11, PPS03, RKS89, ŞSW07, SG09, VSTN06, Zil87]. Romeux [CJ95]. Ron [Gol11b]. Ronneby [Bv97]. Rosario [Bor94]. Roseau [PY05a]. Roskilde [ACL98, AYB^{+}09, CSV05, KRRBV07, SJDT08]. Rostock [FLUV02, HU08, TDG91]. Rotorua [MJR04]. Rotterdam [CPSA10, KvdM00, dS11a]. Rouen [AP93, CMZ99, CL90]. Rouge [ANS^{+}09a, ANS^{+}09b]. Rough [PSGB^{+}04, PS05b, PS06, PS08b, PSR08, PSW^{+}09, PS10, PSS^{+}10b, PSC^{+}11, PSS^{+}11, PS12b, PSD^{+}05, PS05a, PSD^{+}07, PSM^{+}07, ZSO99, APSZ02, ASR^{+}07, CGBZ08, GHH^{+}06a, KPRS07, KŚHM11, LNW^{+}12, PS98b, SCH^{+}09a, SWS^{+}05, SYP^{+}05, SKR^{+}10, TSKGB04, Wan03, WPSY06, WLGB^{+}08, WLP^{+}09, YLW^{+}07, YRWS11, YYS^{+}12, YGL^{+}10, ZY01]. Rousse [LVW05, VWY97, VWY01]. routes [Fre03]. Routines [SBD^{+}76, GBDM77, Smi74]. Routing [Sch98c, SPSV^{+}04, BS94, YD98]. Rovaniemi [OKS02]. Rovereto [Pri03a, Pri03b, PQ05]. Rozenberg [BEKS02]. RP [BP09b, DP11, FLP12a, KP10c]. RPC [BMS96]. RPC-memory [BMS96]. RR [CL08a, HL10, KS12d, MPdSM07, PS09b, RG11]. RRG [EW05]. RS [PBK^{+}10]. RSA [Abe06, Dun12, Fis09, Kia11, Mal08, Men05, Pie10, Poi06, Joy03, Nac01, Oka04, Pre02]. RSCTC [CGBZ08, GHH^{+}06a, SKR^{+}10, YYS^{+}12, APSZ02, PS98b, TSKGB04, ZY01]. RSEISP [KPRS07]. RSFDGrC [ASR^{+}07, KŚHM11, SCH^{+}09a, SYP^{+}05, Wan03]. RSFDGrC’99 [ZSO99]. RSKT [LNW^{+}12, WPSY06, WLGB^{+}08, WLP^{+}09, YLW^{+}07, YRWS11, YGL^{+}10]. RT [SBF08]. RTA [Baa07, Gie05, Pfe06, Tre09, Vor08, Bac00, Boo91, Com97, Der89, Gan96, Hsi95, Kir93, Mid01, NR99, Nie03, Nip98, Tis02, vO04]. RTA-89 [Der89]. RTA-91 [Boo91]. RTA-93 [Kir93]. RTA-95 [Hsi95]. RTA-96 [Gan96]. RTA-97 [Com97]. RTA-98 [Nip98]. RTA-99 [NR99]. RTBI [CLZ^{+}09]. RTCSA [CH04]. RTSE [BR98]. Ruda [Mat01]. Rule [AST05, AB04, CL08a, HL10, KS12d, MPdSM07, PB07, PS09b, RG11, SW03b, ZZ02, BGP08, BGP11, GHP09, OPS11]. Rule-Based [BGP11]. RULEML [AST05, BGP08, BGP11, BG12b, DHRT10, GHP09, OPS11, PB07, SW03b, AB04]. Rules [DHRT10, GB97, SF08a, AST05, AB04, BG12b, SW03b, Sel95]. Run [Dwa00, O’H98, PA01]. Run-Time [Dwa00, O’H98]. Runtime [HNRW06, Bre97, BFF^{+}10, BP09a, KS12b, Leu08, ST07]. RUR [DvV96]. ruSMART [ABK12, BMK08, BMK09, BDK10a, BKH11]. Russia [AM10, AN97, ABK12, BMK08, BMK09, BDK10a, BKH11, BUG93, BBP93, BBP96, BBZ00, BBZ01, BGU94, BGU95, BKLS11, BZ03, CW94b, CVV12, DVV07, FMRW09, GSP01, GPS03, GLS05, GKS05, GZSC07, GHH06b, HRSS08, HKLP12, HM11c, KS10a, KS12c, KHI06, KHS07, KV11, KMKP11, KŚHM11, LKZCS05, Mal95, Mal97b, Mal99, Mal01, Mal03, Mal05, Mal07, Mal09, Mal11, Mar05a, NT92, NM94, PVV10, RDK09, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, Vau06, VV07, Vor92a, Vor92b, Vor93, WPZA11]. Russia-Taiwan [HM11c]. Russian [BMK08, Vor92b]. RV [BFF^{+}10, BP09a, HNRW06, KS12b, Leu08, ST07]. Rydzyna [Dem80]. Rye [KJR05].

S [DKN12, M^{+}06, PPPM10, BEH^{+}85, CW94a, MR97b, RTA95]. S-Cube [PPPM10]. S-GRACE [M^{+}06]. S3 [HR08]. S3E [RKM^{+}12]. Saarbrucken [Loe74, MT97, Meh85, NW01, NDF94, RW86, AL11a, Bar11, BSL^{+}09, DGG^{+}11a, DH11, ENO12, GKS12, GO11, GK12, Hof11, HMMW10, JK00, Kno11, MR12a, MP12, Pat00, PS11b, RLKHL12, SOP12, TSS11]. Saarbruecken [BP05]. Saariselkä [KTH09]. SAB [AHMT08, DGG^{+}10, NBC^{+}06, ŞS05a, ŞSW07, ZBH12]. Sabadell [TN08]. SAC [AMW07, AKS09, BY07, BGS11, HH04, JRSN09, JY14, MV12a, PT06a, MZ04, NH03, ST01, TM99, VY01]. SAC’99 [HA00]. SACH [DJ08]. Sackville [AKS09]. Sad [CIT11, Ong11]. Safe [BGH06a, BBH^{+}05]. SAFECOMP [BRS09, FBV11, Gór06, HS08a, OD12a, OD12b, SO07, Sch10b, WGD05, HLW04, ABM02, AFL03, Ehr98, KvdM00, Vog01]. SAFECOMP’99 [FKP99]. Safety [AFL03, BR12a, BRS09, DG06b, DPV11, DCFY10, FBV11, Gór06, GP09c, HS08a, HLW04, KvdM00, KS08b, MT07a, OD12a, OD12b, PVW10, RS12, SO07, Sch10b, XLKZ12, ABM02, Ehr98, FKP99, Vog01, WGD05, BMU^{+}09]. SAG [HPR05]. SAGA [HKNW07, LKZCS05, WZ09, AM91c, AS03, Ste01]. Sagamore [yF75]. SAGT [KKS10, MP09, MS08b, Per11c, Ser12]. SAIG [Tah00, Tah01]. Saint [APV09, BN09a, BL92, BFGM05, BHH04b, BHH04c, CJ08a, CD07, CCM08, DL04a, FS07d, FCDB07, GP89, GQ95, HR95, KR02b, KJB09, Mar05a, MNW98, PR05, TST09, VMvN12, Coi99]. Saint-Etienne [APV09, TST09]. Saint-Gilles-les-Bains [KR02b]. Saint-Malo [BN09a, BHH04b, BHH04c, CD07, CCM08, DL04a, FS07d, GQ95, KJB09, MNW98, VMvN12, Coi99]. Saint-Michel [BL92, PR05]. Saint-Pierre [GP89]. Salamanca [BvDMST11, BCC07, BFC^{+}09, CSPC09, CWO^{+}09, CSA^{+}12a, CSA^{+}12b, GHP^{+}08, ORB^{+}09]. Salerno [DF09, SVS08]. salesman [Rei94]. Salford [MM04, Wat95]. Salil [Gol11b]. Salomaa [KMPR04]. Salt [FKM08, LLS00, LMT00, SS07a]. Salvador [Agu12, MSAdAS05, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d]. Salzau [ES02]. Salzburg [DKU05, RT07, TOL08, TdRM^{+}09, Wii05b, Zim92, ZVU99]. SAM [GR06b, KH11c, She03]. SAMOS [BDW08, BDSW09, HPTV05, PV04, VWH06, VBH07, APS08, DKK08, KGL04, KLB^{+}06, Lop06, PAD08b, VP04, vdMBD08, DTV02]. Sampling [CN06]. SAMT [AKSO06, CKM^{+}09, DGG^{+}11a, DHH^{+}08, FSA^{+}07]. San [AD05, Abe06, ADJ^{+}01a, ADJ^{+}01b, Ayk09, BC07, BFMS08, Bor93, BDV95, BKM06, BKM08b, BG07b, CS05a, CGP02, Cha02b, CGS07, CMM93, CUdlC04, CRS10, Cou03, DG07b, DH94, DF96, DS03, Duf09, Dun12, EW05, EIRW06, FV04, Fis09, GGD09, GAA06, God05a, Goe98, GZ09, Gup99, Har09, HH03, HW08b, IR08, IB07, Jac09c, Jac09d, Jac09a, Jac09b, Joy03, Kar09, KS12b, Kia11, KMSS02, KPR07, KP15, Kur09, LSW02, LG94, LSgL^{+}06, Lin14, LPZ08, LR05b, LHH11, MWYF10, MT09c, Mal08, Man75, MS00a, MTO99, MRY^{+}11, MNPW07, MZC06, Men05, MS02c, Mue01, MO92b, Nac01, Oka04, OZ09, PS97a, PZ96, Pie10, Poi06, Por98, Pre02, RLL04, RA00, Rei09, Rol99, RCS10, RF99b, SS09a, SPCG07, SEG09, Seh97, Shu09, Sle92, SS09c, Ste09b, Ste09d]. San [Ste09c, SWB03, TD07, VH04, ZSM^{+}09, dVGR04]. Sanda [KKKN05]. Sanibel [FSM03, SW03b]. Sankt [Her91, LK07]. Sans [Gau93]. Sanskrit [HKS09, Jha10, KH09]. Sant [PB97]. Santa [AH02, Arm02, ASAM07, Ban92, Bel00, Bon03, Bri93, Buc90, COT98, Cop95, Des94, DPP95, Dwo06, Ell95, FK87, FR95a, Fer96, Fra04, GT04, GF12b, Hal09, HT06a, ID03, ID06, Kal97, Kil01, Kob96, Kra98, MS10a, Men07, NNPW12, Pal00b, Rab10, RMDS10, Rog11, SNC12b, ŞS05a, San92, Sho05, SSM08, Sti94, Wag08, Wie99, Yun02]. Santander [GdlP99, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e]. Santiago [HN12, LFL05, MOBS06, MR07a, MFÁ^{+}09a, MFÁ^{+}09b, MOTN07, RBTD04, VRHJ09a, VRHJ09b, ZBY07]. Santo [BCM08]. Santorini [CV08, CD01, CDND11, GVT08, Had03, KMMT12, MPK04, Nik05, NCJK08, ONPR07, PT09, SB12a]. Santorini/Thera [CD01]. Santos [dCVK07]. Sanya [DD06, LTW11, WDGL10a, WDGL10b, WGLL10]. Sao [BC10b, CG02, Sim92, BL04a, dRCVT11, SVA05, SM11b, dLWC03]. SAPF [vdL01]. SAPIR [DLdS04]. Sapporo [AH07, GJT03, GTY03, HO03, LSL^{+}09, WZ09]. SARA [MR07b, ZS05b, CW00, KH02]. Saragossa [Mau96]. Saratoga [HMRT05, Kap92, Mur99]. Sarawak [CFL10, RWSS12]. Sardinia [AGL06, HH02]. SAS [AV08, CS96, CM11, Gia04, HS05a, MS12c, Myc95, NF07, PS09a, Yah11, Yi06, Cou01, Cou03, HP02b, Le 94, Pal00b]. SAS’97 [Van97a]. SAS’98 [Lev98]. SAS’99 [CF99]. Sassur [OD12b]. SAT [BW05a, BG06, BZ08b, CS12a, HM05a, Kul09, MSS07, SS11a, SS10, GT04]. Satellite [BW03a, EIRW06, IB07, Ölv10, PJC^{+}12, Dur12, IWC04, JSdSA05, Bru06b]. Satisfaction [BJN03, JNS05b]. Satisfiability [BW05a, BG06, BZ08b, CS12a, GT04, HM05a, Kul09, MSS07, SS11a, SS10].

Satoru [TJHS94]. Savador [ZdC08]. Savannah [GS09b, JMO09]. SAWSDL [MHD09]. SBIA [B^{+}12a, BL04a, dRCVT11, SCR06, ZdC08, BR02, BK96, MS00c, WC95, dO98]. SBLP [dCJB12]. SBMF [DSS11, GN12b, OW09, SM11b]. SBP [CSM10, SYNC11, YGE12]. SC-2 [RA00]. SC-3 [RTB02]. Scalability [FFM05, SP09]. Scalable [Dwa00, O’H98, PA01, Paz99, Sto04, WR01, BG11a, DH10a, GP09b, GL08, HR99a, HLFS12, PS07b]. Scale [AMZ09, Dad01, DKK08, DFP^{+}12, DKW07, GLZ^{+}03, GL03, HKW09, HKW^{+}10, HKW11a, HKW^{+}11b, HKW12a, HKW^{+}12b, Ker01, KSW05, Nie99, Pol01, SvdMOP06, TO08, ter97, Kåg98, KvG98, BtHRBB12, CG12, GL03, Ker01, KSW05, LMWY04, LMW08, LMW10, LMW12, Nie99, SMP07, TMLL09, TO08, ZH00, ter97, LMW06, MWY01]. Scale-Space [Ker01, KSW05, ter97, Ker01, Nie99, ter97]. Scale-Space’99 [Nie99]. Scaling [MG08]. Scandinavian [AF06, AI98, BG03b, EP07, FK12b, GK90, Gud08, HK04, Hal00, HK11, KPK05, Kap10, KL88, KL96, NU92, PS02b, SHJ09, SS94]. Scarborough [DD07]. scattered [Vel94b]. Scenarios [Fer08, LS05]. Scene [DFP^{+}12]. Schaumburg [HBS00]. Scheduling [FR00, FRS03, FS07d, FS08, FS09b, FS10, Nar01a, FR95a, FR96, FR97, FR98, FR01a, FRS02, FRS05, FFRS05, Lis89, Nar01b, RF99b]. Schema [Sch03b, BdBC01]. schemes [Eng74, Gre75, Pfi96]. Schloss [BHP95, Fre98]. School [ACP01, ABDW10, AAB^{+}07, ABW10, BGHJ07, BBF^{+}06, BBM^{+}08, BDPS02, BI03, BH07, BDZ08, BPZ09, BI11, BCP12, Bjo80, BCSW10, BC08, BBPP09, BHK01, Cas01, CSW06, Cho88, CB06, CJ95, CMT01, CFR92, CCR86, EM05, EK12, ÉEMS04, EHMM09, ECV^{+}10, EEM^{+}11, FLVS11, Fil91, GLW07, Gib12, GG98, GP89, GI03, Gue90, HMT99, Hor06, HPSZ08, HPZ10, JM95, KPS09, LSV06, LVS08, LMŠT01, MSY08, MS03c, MN12b, Mül10, OS09, Pin89, Pod95, PdA^{+}11, PvS07, RUL04, RRU05, RG03, RV98, SHO99, TFE^{+}09, Vää99, Zil87, ZHP12, dBdRR89, dBdRR91, AM91c, BCG02, BS76, BC04a, BB05b, BJ86a, JJ03, LMS96, MST92, MSS05b, Paz97, TBG05, VU05, dBdRR94, BB79]. school/symposium [dBdRR94]. School/Workshop [dBdRR89, dBdRR91]. Schools [DF09, HKV10, KM11, MN12a, Mit05, Mit06, MS08a, BvLR04]. Schwerin [Ham91]. SCI [HR99a]. SCI-Scalable [HR99a]. SCIA [EP07, HK11, KPK05, SHJ09, BG03b]. SCIE [Paz97]. SCIE-97 [Paz97]. Science [ÅBE12, AM10, AI06, ABH12, AS02a, AvAS06, AvASD06a, AvASD06b, AvASD06c, AF02, AH03, Ama08a, AM98, AM00, AJRS03, AN07, AN09, AKG06, AP07, ADR08, BM03a, BBdB12, BMS04a, BMS08b, BJM09, BW10b, BFG^{+}12, Bir12, BZ83, BL03, Bod03, BI04, BBR88, BBR89, BBR90, Bra02, BVNW87, BW00, BKM07, BHH^{+}06, BHK01, BEFP08, BvASD04a, BvASD04b, BvASD04c, BvASD04d, BvADS08a, BvADS08b, CGMQ08, CPRS01, CDV03, CRS11, ČGH^{+}11, Cer07, Chi95, CL90, CJ95, CW88, CKC11, CTS07, CMB^{+}08, CF10, CEER96, Dat09, DV10, DDHP05, DPC^{+}07, DH04, DVV07, DR02b, DKN12, DH07b, DT06b, EM02, EHM11, EFW93, EMW94, Ési06, FRvKS10, FR01b, FKK04, FJ92, Fom06, FM84, FMRW09, GCJB09, GLP12, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, GT08, GTW^{+}08, GT09a, GTM09a, GTWC09]. Science [GT09b, GT10, GTA10, GTM10a, GTM10b, GT11a, GTSS11, GT11b, GTM11, GTP12, GT12, GKB^{+}08, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, GG07a, GG07b, GG07c, GML^{+}08b, GML^{+}08c, GTM^{+}09b, GTM^{+}09c, GM06a, GJRT01, GRS06, GSLM12, GS10, GK09, GTY03, GHH06b, Gue90, HMV01, HLKX12, HRSS08, HKLP12, HK10b, HJRS99, Hof11, HNW04, HL00, JY96, JSV11, JS01, JFBH08, KG12, KM08a, KL95, KP00, KJ09, hKA10, KSW03, KK11d, KBČ^{+}12, KU06, KN09b, Kuc̆02, KK07, KWT03, KR11, KV11, KLT09, LGK^{+}04c, LGK^{+}04a, LGK^{+}04b, LGK^{+}04d, LLW06, LSS02, LYhL^{+}01, LM04, LN04b, LRS98, LDC^{+}09, MWD06, MB82, Mah04, MT04a, MP95a, Meh85, MT99a, MVN86, MC89, MMK98, MMH07, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MGM^{+}12a, MGM^{+}12b, MGM^{+}12c, MGM^{+}12d, MS11b]. Science [NR00, NKM^{+}09, OSR^{+}02, OT08, OS07, Ong05, Ong10, PAD^{+}08a, PDML11, PR03a, PS12a, PH10, PRK12, PHH10, Pin89, PR96, RRR99, RS06, RMFG06, RT00, RM97b, RV03, RSW12a, Sar03, Sas05, Sch06, Sei07, Sha86, SdDK91, SvADS07a, SvADS07b, SvADS07c, SvADS07d, STDH02c, STDH02a, STDH02b, TGM^{+}10a, TGM^{+}10b, TGM^{+}10c, TGM^{+}10d, Thi10, TW07, TLJ06, VPBŠ03, Vet11, Wal04b, WTP^{+}06, WZA10, XKGS12, XL11, ZS07, ZZZL12, ZJS^{+}02, dA09, dS02, dS11a, vLIvdH^{+}07, vLMP^{+}10, vdMvdT08, AN97, ADJ^{+}01a, ADJ^{+}01b, ANS^{+}09a, ANS^{+}09b, ABD04, AF99b, AS02b, AR98, BBEdK98, BB91, Bec75, Bec79, BV92, BS92, BMP90, BN91b, Bli75, Boh75, Bor91, Bor92, Bor93, BGM94, BS93, BL01, BFJV97, BvADS08c, Bun96, CV96, Cho88, CK84, CJK88, CER79, CS00b, CLP05, CK82, DK90, Dem80]. science [Deu81, DEM96, DF96, DD05, DMWW88, DMW96, EFM04, ENR83, Ehr87, Ehr89, EKR91, EK81a, FHRR05, FR99a, FM99, Fri01, Gor03, Got83, GAY86, GS88, GGS99, GP89, GSV05, Gru77, GC81, GRW86, HK92, HS00, HM01, HS98c, HO98, JM02, JS05, JS84, KSHS01, KK87, KSS02, KS05a, Kra05, KMS89, KKGT03c, KKGT03a, KKGT03b, KPW99, Mad89, Mah85, Mau91, May93, MST95, Maz76, MT89, MR97b, Moh91, Moh97, Mon01, MSH05, MRS97, MO92b, Nag90, Nag95, NDFMS97, NT92, NM94, NT98, NE02, Niv98, Nol81, Nor86, Nor87, NK88, Nor90, PT95, PDHS03, PS96a, PPR87, Pit89, Pit91, PRJ95, PRR94, PR97, RS97a, RRR01, Rol97, RV02, Rov90, SKMR94, SE94, SPK01, Shy92, Shy93, SU97, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, SvASD05a, SvASD05b]. science [SvASD05c, SA04, Tar91, Thi94, Thi95, Tho99, TS87, Tiu00, VBCBS05, Wei79, WNE99, WH95, Win78, YT99, ZHF04, ZBG98, vB97, vL89, vL94b, vL95, vLWH^{+}00, AS02b, TWW77]. science-ASIAN [SU97]. Sciences [ADK04, BCBF08, BGF06, BR12b, CBT07, CR05b, DMWW88, GL74a, GL74b, LK10, LNE06, LSY^{+}06, LR05b, PMH09, Rah04, SXC04, Squ09, YŞ03, YGGÖ05, BGD^{+}05, DT07]. Scientific [BSW12, Cai05, DMW06, DT07, FHJ^{+}02, GMV06, GMV07, GMV09a, GKMV10, GKMV11, GKMV12, Jón12a, Jón12b, KEDW07, KS07a, LMWY04, LMW08, LMW10, LMW12, MFÁ^{+}09b, PF07b, Stu90, BZ83, Des86, DW94, GMV05, Jah93, Kåg98, LMW06, Lub95, MWY01, Mic90, MT86, PD96, RKS89, TU94, WL03, AB12, CFB11, GL10b, GAR03, GAR04, GRR05, HPR05, Ish97, LM08b, Win09]. Scientists [DK90]. SCM [Con97, Mag98, Som96, Est95, Est99, WvdH03]. SCM-4 [Est95]. SCM-5 [Est95]. SCM-6 [Som96]. SCM-7 [Con97]. SCM-8 [Mag98]. SCM-9 [Est99]. SCN [BC05, DY06a, GD10, ODV08, VD12, CGP03]. SCOPES [Kra03, Sch04a]. Scotia [Deh97, Tom92]. Scotland [AFS02, ACH00, ABM05, BJS99, BJ01b, Bra02, Cag00, CG07, CHD98, Coh10, CP02, Dyc00, ER05, FRS02, GL92, Joh01, Jon00, Koz04, MM01, Mil00, MK96, Pol00, RMS04, ST88, ZS05b]. Scottsdale [KRMM^{+}02]. SCVMA [Mac06]. SDC [DFF06]. SDKB [ST08, ST11]. SDL [GNR07, OO12, PRR05, RR01, RBG09, GNR07, OO12, PRR05, RR01, RR03, RBG09, She03]. SDM [JP04b, JP05b, JP06, JP07b, JP08, JP09b, JP10, JP11b, JP12a]. SE [TC95]. SE-HCI [TC95]. SEA [Fes10, Kla12, PR11, Vah09]. SEAFOOD [MJ07]. SEAL [BOH^{+}12, DBC^{+}10, LKZ^{+}08, WLC^{+}06, McK99, YKF97]. Search [BGS^{+}03, Nar01a, NP12b, SVS08, SBH07, SBH09, Hof03a, Nar01b, Wal99b, CB10, CB11, CB12, CÓ11, FdS12, Fri06]. Searches [Yu02]. Searching [TK05]. Seaside [OPKR05]. Seattle [BBMB^{+}08, BJ06, BG06, BS94, DSB04, DPB^{+}04a, DPB^{+}04b, DSM03, ET06, FRS03, FS08, FS06, GA09, HNRW06, JP11b, Luk97, MW09, MT02b, MNLM06, NP12a, Pfe06, VZ06]. Sebastian [CGP02, CUdlC04, DF96, CRS10, GGD09, RCS10, SPCG07]. SeBGIS [MTH05, MTH06a, MTH06b]. SeCoGIS [CVLL12, DMB^{+}11, HRK^{+}07, HP09, SPC^{+}08, TDK^{+}10]. Second [AP93, AG98, AS03, AC09, AFS02, AHK07, AFGvE11, AH98, ABH^{+}10, AF99b, ABD03, Auc98, AKTT12, Ayk07a, Ayk07b, AKR^{+}09, BBCL08, BA06, BHL^{+}03, BMK09, BST03, BBF^{+}06, BKA11, Bar08a, BP12a, BBH^{+}10, BGH06a, BR04, BMP01, BBP^{+}06a, BBP^{+}06b, BH08, BBLT06, BS06b, BMR^{+}04, Ber98, BCV06, BGF06, Ber03, BB01b, Bha02, BZU07, BBP96, BBRS05, BL06b, BKLP11, BMN05, BMW10, BO00b, BFJM09, BCSW10, BDDS05, BHJ^{+}11, Bou11, BCRT92, BFMR96a, BFMR96b, BSB^{+}99, BR07b, BRT11, BKM08b, BD08, BLPW02, BC98, BFOY04, BJ01d, But10, CC95, CL08a, CKP^{+}12, CSJD06, COT98, CGS01, CD09, CL07, aCF06, CZ05, Cha02b, CLL^{+}09, CMZL04, CDLZ06, CMM09b, CY11a, CP06b, Chi04, CJ06, CKO01, Chr00b, Chu03, CFVV^{+}12, CFVCF^{+}12, Coe02, Coi99, Col07, CFR92, CMM01, CRW^{+}12]. Second [DYCD09, DF01a, DSLT10, DM01b, DKS^{+}07, DKM91, DPW04, Des09, DEQ92, DVV07, Dik04, DS03, DO09, DP97, DHMM96, Duf09, DKN02, DFGN12, DKW07, EDGP11, EM02, EGO09, ES02, Ekl04, ERK07, ET01b, EW04, EW03, ECV^{+}10, FSA^{+}07, FV03, FMC04, FSM03, FM04, FMW01, Fin99, Fon03, FRL07, FP08, FGK^{+}11, FR99b, FGK^{+}10, FGM92, FKM08, Fre96, FK09, FM96, FMY04, GVO^{+}10, Gae96, GKL09, GP08a, GWY99, GANACBR10, GLCT10, GMV03, Gel01, GK06, GDP07, GAA06, GO09, GSYZ10, GSH^{+}04, GPE05, GW98, GPS03, GZSC07, GL08, GCC^{+}02b, GSS00, GH93, GS06, GG02b, HT09, HRdV11, HP99b, HKM95, HS08b, HM12b, HKS06, HHP09, HNM02, HS02, Heu00, HRT^{+}03, HRR^{+}06, Hir98, HJRS99, HCR06, Hor00, HPSZ08, HM11c, HKS09, cHPkC^{+}07, HK07, HU05]. Second [IT06, IMK06, IZ99, IZP09, JY96, JYZ04, Jen03, JPD04, JWbG^{+}06, JWG^{+}06, JP05b, KS03a, KCR04, KMT00, KKGK11, KLZ07a, KA12, KUA00, KMF06, KMS^{+}11, KS12b, Kil05, hKhLKŚ10, KR01, KW98, KP01b, KFLS07, KM00, KYT01, KMN94, Kur11, Lac10, LRS03, LLMT04, LSW00, LP97, LLP99, Lee01, LYMM05, LSgL^{+}06, LMMN10, LOTY05, LO98, LwCM00, LWM^{+}08, Li96, LSDN04a, LSDN04b, LZhL06, LYL06, LCB97, LC11, Loe74, LK09b, LPK05, LH08, LRS98, LSW08, LR05b, LDC^{+}09, MOP11, MKPV05, Mag03, MMMP12, MSG07, MNS05, MZ99, MAD^{+}04, MBW08, MPB05, MEAC04, MdRIR08, MdlBP05a, MdlBP05b, Mar03, MF01, MVOF08, MTCOK10, MWZ10, MKT98, MMOS99, MZCC09, MP09, MSAdAS05, McK99, MRSV07, MR07a, MNW08, MÁ07a, MÁ07b, MKR07, MSB07, MBF08, MMH07, MGSM12, MNTY05, Myc95, NDB94, iNR08, NP11a]. Second [NT92, Nes95, NJHJ08, NLŚ10a, NLŚ10b, Nie99, NS98, NR06, NLS08, NG10, OB07, OS04, OPT01, OSC^{+}08, Pai09, PPP07, PCN10a, PCN10b, PCN10c, PSDS11a, PSDS11b, PMMR06, PvES12, PH02, Per01, PBGK02, PNB04, PS03, PMD92, PBK03, Pol99a, Pol01, Pol99b, PC00, Pre09, PWY08, RSV07, RD01, RW06, RMS^{+}05, RMB10, RR04b, RF12, RBSL06, RH98, RS09d, SW07, ŞSW07, SL11, San92, SVS02, SS01, SW03b, Sch07, SSN08, SM06, SW08, Shu07, SMK01, SMPU98, Smi01, SZ00, SK08, Son00, SMK06, Squ09, SF08c, SS06d, SSB90, SSKR02, Str99, SBH09, SWSY11, Tah01, TMLL09, TMG10, TLF01, TSCW11a, TSCW11b, TWYhL01, TC99, TT10, TY11, TDS94, Tho99, TG00, TBH03, TC98, TNM05, Tou96, Tra04a, TST09, Tuc01, UBM02, VvS99, VM99]. Second [Vck99, Vol93, Vor92b, VWY01, Vyt91, WK08, Wan04, WLY05a, WLY05b, WLY05c, WJ05, WFZL09, WW02, WPM06, Wii05a, WL03, WY98, WWC02, WKK98, XXX^{+}09, Yak02, YD98, YJ04, YDW08, YLW^{+}07, YLS^{+}12, Yas00, YC11, ZRP^{+}11a, ZRP^{+}11b, ZBRL12, ZGK^{+}07, ZS07, ZCDT10, ZS10, ZZZL12, ZG03, ZNCF03, ZY01, ZSP10, ZQ98, dBBG04, dBBC02, dKIM^{+}07, vEGSW11, vdBGG10, vdL98, BYGP95, BDGM05, BM05b, BR06, BMW93, BCP94, Bor94, BJS93, BTF05, yCW96, CYGL04, CO94, Cha05a, Coh96, DWSK05, Der92, Des86, DCP95, DMW96, DLW06, Ehr04, EHSH92, FM97, FR95b, GL97, Her92b, HSS96, HW05, Hut96, JNS05b, JK05a, KC05, KW90, KK95, LMS96, LDTU04, LMV98, Luo05, MS96a, MW03, MTW05, MSS05a, MZF94, Nak92, NcC99, NOOS05, PTRH05, PS98a, PSO05, Pre95, Sel95, SLC01]. second [Ste94, TvdBI02, TYNM05, Van05, VPM97, Vit95, VT97, VS02, YZZ^{+}05, ZL97, ZGT05, Zip92, vL94a]. Secondary [HKV10, Mit05, Mit06, MS08a]. SECUBIQ [DsSL^{+}07, EYX^{+}05, ZSY^{+}06]. Secure [AJN12, BGH06a, BLS12, Bau99, EWZ11, JMK09, LPSW04, Lau12, Lin03, MRZ09, MWZ10, OS07, SH12b, BBH^{+}05, BP95, BY06, CZZ08, JP04b, JP05b, JP06, JP07b, JP08, JP09b, JP10, JP11b, JP12a, JC12, VJ99]. Security [APFV09, AB10, ARS03, Abe10, AGM05, AG07, ABG09b, AG11, ABB10b, And94, AFL03, AZ11, AL11b, APP12, Atl08, AD11, AJN12, BN09a, BBCL08, BA06, BB07, BLO^{+}07, BLW09, BYLJ11, BW11, BSZ12, BKA11, BA07, BMU^{+}09, BDE12, BS02a, BSN06, BGKY08, BL07a, BDC12, BFMS08, BFJM09, Boy01, BM03c, BN09b, BC11b, BTMI11, BRS09, CKD11, CK12a, Cer07, CY03, CWZ^{+}11a, CWYC12, CMZL04, CWY^{+}06, CDLZ06, CRW08, CMS08a, CYCL09, CChL^{+}10, CY12, CCBW11, CSN14, CSY09, CY11b, CKOS09, CBCGA12, DL06a, DDS12, Dan12, DG06b, Dat09, DFR02, DW07a, DSB10, DLNU11, DC12, DPV11, DGM09, DV09, DG10, DEG11, DG12, DQBZ02, DBPZ05, Des09, DR06, DD07, DGDM^{+}11, DMRS06, DMRS07, DG09b, DCFY10, EKS10, Feh11, FLY06, FHLP09, FHKQ12, FBV11, FG01, FG04a, FJ10, FYM12]. Security [FHW08, FKL08, FLP11, FMY04, GKL09, GLMP07, GMO09, GANACBR10, GANACL11, GANACBdCdV12, GAL12, GTM09a, GTM10a, GTM10b, GMS06, GF12a, GCM08, GPS03, Gór06, GPT10, GW12, GV09, GP09c, GJTS06, HY12, HS08a, HS08b, HLW04, HWG10, HK10a, HC11, HN12, IKY05, IN11, JM05, JL08, JM11a, JYZ04, JM11b, JP12b, JY12, KMR^{+}05, KLB^{+}06, KLS10, KY07a, Ker12, KM96, KYL07, Kim12, KvdM00, KS10a, KS12c, KS08b, Kur07, Kur10, KDWW10, LC06, LYL11, LZL11, Lai03, LOX99, LPT07, Lau12, Lee04, LYY07, LC09, LH10, LW11, LY12, LM06b, Li11b, LL04, LTW11, LM03, LYL06, LY01, LPK05, Lop06, LH08, LT11b, Mac12, MBH^{+}09, MT07a, Mat09, MF08, MG07a, MZC06, MKR07, MP12, MSS08, Mue06, Nac12, NR07, NQL06, OiDP98, Oka00, OSC^{+}08, OALZ^{+}08, OD12a]. Security [OD12b, PH11a, PCA^{+}09, PYLW08, POS00, PGD07, Pie08, PZ09, PMC06, PG09, PR98a, Pri03b, QOZ01, QGZ03, QIW07, QMW09, QSWL11, QBY^{+}12, RL06, RN11, RS12, RB10, Roy05, RSW12b, Sad13, SP09, SNČ^{+}12a, SNS03, SN08, SO07, SRG04, SYMA09, STP^{+}10, SMBQ07, Sch10b, SSN08, SP08, SG09, Shi06, Shi07, Shi09, Shi10, SEK^{+}11, Shi12, SK12, Sio10, SCD^{+}10, Smi12, SG03, SQL10, SH10, SD10, SLM07, SMS12, TM09, TWQ^{+}12, TPCB12, TQYX11, Tsu08, VG12, WPV04, WmCL07, WS12, WK06, WLRL08, XW11, XLKZ12, XBM12, Yal10, YZC^{+}07, YCC^{+}08, YSR^{+}06, YY09, YLL09, ZGK^{+}07, ZZ04a, Zhe02, ZYH03, ZYB06, ZY10, ABM02, AJ01a, AJ01b, Ber96, BB89, BCD00, BG05, CG02, Che03a, Cup00, Dam99a, DMO98, DF01b, DB98a, DWML05, DEQ92, DKU05, Ehr98, FBKS03, FKP99, FH01]. security [Gol94, GKW02, GSP01, GKS05, HOQ97, H^{+}05, JY96, JW05, Kim02, LL03, LY05, LQO04, MZ99, Oka03, PC05, PY05a, PS96b, PSNS99, PGV93, QMLW05, Qui98, Son00, SKY06, VPM97, VM99, VM01, VJ99, Vog01, WGD05, Won01, ZLDB05, dSG05, BC05, BKL^{+}12c, BGW06, CHPW05, CCLR98, Chr99, Chr00a, Chr01, CCMR02, CCMR03, CCMR05, CCMR06, CCMR07, CCMR09, CCMR10, CMMR11, CCMS11, CMSA12, CGP03, CPPB06, CLBP11, DY06a, GD10, HMSU04, HU05, JMT11, Lom97, MFG12, MTW05, ODV08, San02, Sch03c, SK12, SMCM07, Vig98, VD12]. SeDeS [MDH10c, MDH11a]. SEE [MRY^{+}11]. Seeon [ABB99, SSU12]. SEFM [BPS11, EHH12]. segmentation [NMS93]. SEI [For88, Gib89, Dei90, DH94, Ibr95, Sle92, Tom91]. Seine [Pod95]. SEKE’99 [RB00]. Selangor [ZRP^{+}11a, ZRP^{+}11b]. select [FNY93]. Selected [AS11a, ABG^{+}06, AYB95, Ada96, ABH12, AZ96, AGP08, AS99, ABF^{+}07, AKT06, ACK^{+}07, ABKT11, ADB^{+}12a, ADB^{+}12b, ABMS06, AV07, All96, ABdJS10, AKR^{+}12, Alp11, ANR99, AM07, AC09, Ama08b, APW06, AS98, AGT98, AMNT06, Apt00, AFR^{+}04, AS10, AÖ12, AS12, ABGL09, AK12, AS07b, AL11b, AOSV08, APV09, AM02, AAR05, AMM99, AHSC11, ACHH12, AJN12, AKS09, ABRS06, ABFR07, BBMB^{+}08, BBD^{+}10, BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, BdBC01, BFGM05, BYLJ11, BL12, BN94, BKLM07, BHKR11, BHJKPG06, BGH06a, BF08, BDE12, BB91, BUG93, BPR05, BBP^{+}12a, BBP^{+}12b, BM11a, BDG12a, BDG^{+}12b, BP09a, BC96a, BCD04, BDd09, BHM^{+}04, BHMS07, BMS04a, BCM00, BCV06, BB01b, BLSS11, BG12a, BM99b, BC93, BVAT12, BY07, BZU07, BMS95, BGS11, BBH01, BS10, BJN03, BPT06]. Selected [BRST08, BK02, BB10, BGU94, BGU95, BDC12, BFSN10, BPD06, BCC^{+}12, BMN05, BLvdH10, BB06, BFMS08, BFJM09, BDDS05, Bor93, BGM94, BG07a, BCC07, BGL09, BGJ11, Bos00b, BFT^{+}08, BDN08, BRM06, BKL^{+}12c, BGKS04, BR07b, Bra07, BC11c, BBT10, BB11, BGHS05, BBK^{+}11, BAdVR08, BHLvdP07, Bri95, BGY96, BKM06, BKM08b, BtHRBB12, BSH06, Bru06b, BW09, BS12, Bru04, Bun96, BR96b, BC98, BE01, BD03, BR07c, BY06, BTMI11, BDLMM12, BFOY04, BH06d, BCDN07, BGH06b, But10, BGW06, CS00a, CJ11, CG12, CLMP02, CCS04, CC95, CKP^{+}12, CCJS07, CKD11, CL11, CK12a, CV04a, Can12, CGL^{+}09, CBG^{+}10, CBS^{+}12, CHB^{+}12, CV08, CMSM^{+}10, CP12, CS05a, CPSA10, CGMP11, CW94a, CM95, CA08, CR01a, CAS^{+}09b, CW11, Cha09, CWiM^{+}09, CPZ07, CLZ^{+}09, CF09a, CY10]. Selected [CY11a, CYZ12, CHG^{+}07, CD00b, CM09a, CHZ08, CBS^{+}11, CW94b, CEFZM05, CK06c, CS10, CFG96, CBN99, CCMR05, CCMR06, CCMR07, CCMR09, CCMR10, CMMR11, CCMS11, CMSA12, CSN14, CCG08, CSY09, CY11b, CFVV^{+}12, CFVCF^{+}12, CHD98, CV10, CVV12, Coe11, CF09b, Col02, CML^{+}10, CDN12, CUdlC04, CMCS11, CDNS00, CFP^{+}09, CM09c, CKTW12, CF10, CS12c, CvRVSN12, CRW^{+}12, CLBP11, CEER96, DLFP02, DFL^{+}10, DKY06, Dam97, DYCD09, DGT10c, DM06, DG06a, DDS12, Dan12, DF10, DGT96, DDE04, DSLT08, DSLT10, DSHL11, DS09a, DSS11, DDHP05, DPC^{+}07, De 07a, De 10b, DFPV11, DS09b, DHtM^{+}12, DGG^{+}11a, DGM09, DV09, DG10, DEG11, DFP^{+}12, DBB12a, DLW12, Des09, DJNR06, DLN10, DGSN11, DKN^{+}11, DEM96, DR06, DR09, DP10a, DGWB09, DPA07, DD07, DvEF06, DGR06]. Selected [DGDM^{+}11, DMRS06, DMRS07, DS11b, DG09b, DKN12, DPP95, DPP97, DPP98, DFN10, DJ08, Dol09, DO11, DS11c, DFT09, DP97, DMW06, DHMM96, DW06, DW07b, Dun09, DS06, Dwa00, DS07, EJM10, EDGP11, EL07, ELS12, EKF12, EO94, EEKR00, EKP^{+}07, EPVW10, EP06, ENO12, Esc10, EIRW06, EFZKM07, EV09, EHMM09, ECV^{+}10, EVV^{+}11, EEM^{+}11, Est95, EH12, FRS08, FBKS03, FBSMS08, FPFR05, FPRAW04, FS07a, FZJE^{+}05, FZJE^{+}06, Fek08, FKK09, FGB11, FMZ05, FL09, Fia99, FS07c, FKM09, FdCLJ^{+}11, FdCLN^{+}11, FPCK11, FST09, Fle99, Fog94, Fog95, Fog96, Fon00, FS07d, FS10, FKM08, Fre98, FN12a, FPRS06, FLMK06, FLT07, FKLT08, FCDB07, Fur95, FU96, GVO^{+}10, Gae96, GKL09, GML^{+}08a, GPCL10, GANACBR10, GANACL11, GANACBdCdV12, GAL12, GCFA12, GVCA07, GY08]. Selected [GS01, GLV09, GKS^{+}07, GMP06, GW03, GKT09, GKT10, GKST11, GKS12, GHP^{+}11, GPR^{+}12, Gho10, GCK06, Gie08, GKL^{+}11, GH05, GH12a, GPM98, GHSW04, GLM^{+}08, GT04, GOZ05, GGS11, GSS09, GSLM12, GM99, GR06b, GW06b, GO06, GH93, GKMN10, GHI^{+}06, GVT^{+}11, GS06, GB07, GHP^{+}08, GDG11, HS12a, HM11b, Hag02, HDC99, Han09, HLR^{+}98, HLC^{+}12, Har01, HK12, HKM95, HKI^{+}09, HNRW06, HOD96, Hee94, HDBM11, HBH08, Hig05, Hil06, HPS09, HR08, HCF06, Hof97, HhKRT06, HŽ07, HI10, HPRS06, HM05a, HNP12, Hor06, HZB07, HPSZ08, HPZ10, HM11c, HNPP07, HGBF12, HGGG12, HGPH12, HMBD09, HKS09, HHG11, HdMKS12, HKM^{+}11, IM10, IB07, IPSK04, IMW04, IS11a, IS11b, IST07, IGN05, IFV07, IHX^{+}08, IDMY08a, IDMY08b, IKKU08, IS12c, JRSN09, JSS08, JLST06, JLM00, JP11a]. Selected [Jón12a, Jón12b, JDMB10, Jou11, JP12b, JY12, KEDW07, Kah03, KG04, KN09a, KKS11a, KKS11b, KA12, KT96, KS09b, KM08b, Ker12, KS12b, KL12, Kie12, KYL07, KKH09, KSB11, Kim12, KKS11c, KKS11d, KKS11e, KKS11f, Kin08, Klu97, KH11a, KH11b, KTB09, KBHSW06, KHSM^{+}08, KC00, KW10, KKC09a, KKC09b, KS07a, KS07b, KBČ^{+}12, KMS12, KLT^{+}07, KH11c, KBC^{+}09, Küh07, Kur10, KŽ10b, Kut12a, Kut12b, KCK08, LSJ06, LJS08, LV12, LYL11, Lam96, LM00, LS10a, LO11, LS12a, Lau01, Lau12, Lau93b, LLP99, LT11a, wLN99, LYY07, LC09, LH10, LY12, LMSS07, LOST04, LT05, LOTY05, Lei95, Leu08, LS05, LM08a, LWM^{+}08, Leu03, LLP^{+}07, LOS05, LLM^{+}08, LCF^{+}04, LM96, LAF^{+}10, LMP04, LMCH12, LYCV08, LMW12, LL06, LLO08, LW12, LK04, LK07, LK09b, LOH05, LSU07, LMŠT01, LP08, LGS09]. Selected [LSW08, LS09, LV07b, LCY^{+}11, MWD06, Mac12, MMMP12, MP06, MSG07, MG08, MBS11, MBW08, MEK03, MY06, MMBND07, MSRP98, MS06a, MVW09, MŠKL02, MOBS06, Mar11c, MRU12, MS05, MP10b, MR12a, MTV09, MPT10, MPR07, MRY^{+}11, MRPM10, MRP11, MPR12, MMM10, McK99, MFG12, MLTC11, MMG10, MCSN10, Mey95, MS99a, MNW08, MW08, MB09, MSH08, MV12a, MA05c, MNLM06, MP12, MOT12, MK01, MTC^{+}08, MSB07, MvL10, Moo01, MS11a, MF06, MG07b, MDP04, MDPQA09, MDPQA12a, MDPQA12b, MBS05b, MBJ^{+}07, MK12, Mug97, MOTN07, MTNL12, Mül07b, MGSM12, MNTY05, NMM10, NP09, NP11a, NP12a, NZZ11, Neu04, Ngu06, NT98, NR06, NLS08, NS99, NVSB^{+}07, NvRSS11, ND04, Nyb08, O’H98, OROD07, O’S03b, OFR09, OLL10, OS07, OBS12, OW09, Ölv10]. Selected [OPP04, OSV11, OBM11, OK96, OSRS07, PT95, PAV^{+}10, PM08, PHZ11, PVW10, PR07, PT09, PJC^{+}12, PAD^{+}08a, PDML11, PPCS12, PH04, Par98, PMMR06, PSPH^{+}10a, PSPH^{+}10b, PS12a, PPJRN^{+}10, PYLW08, PW09, PvES12, PP12b, PRT06, PS07a, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, PNPP12, Pet02, PNB04, Pod95, PBRB08, PT06a, PQ05, Pro11, Pue07, PS09c, QLXL07, QCDM06, RPR08, RM09, RM97a, RS99a, RLL04, RUL04, RRU05, RTZ08, RKM^{+}12, RTA95, RBF06, Ret97, RSS06, RN11, Ria08, Ria09, RtTMP10, RtTMP11, RtTM12, RK09, RL11a, Rob06, RBSL06, SST00, Sad13, SS08a, SNY06, ŞSW07, SM11a, SSVW12, SFS11, SHN07, SÇS12, SH12a, SINK08, SGG06, Sav06, Sch10a, ST08, ST11, SC11, SJDT08, SNRG11, SSN08, SNZ08, ST12b, SHW09]. Selected [SCL^{+}06, SLL^{+}07, SYY^{+}08, SPÖ^{+}10, SKPG12, SPON08, SM11b, Sio10, SCD^{+}10, SA12, SHP^{+}05, Smi05, Smi06, ST07, SSM08, SCZ10, SOP12, Som96, SMF^{+}11, SKK04, Squ09, SKCK07, SS04, STWS07, SG07, SBF08, Stü09, SV01, SM09, SHŚ^{+}07, SSZ12, TT10, TBC07, TNA^{+}10, TKSN11, TPCB12, TC98, TT06b, TST09, Tsu08, TtHVS06, TWSN07, TNGK08, UWB^{+}06, UBW06, VS09, VAR12, VMvN12, VEG^{+}11, VFC08, VU09, Vet11, Vid12, VSMS12, VS05, VKG12, WK08, Wan98, WmCL07, WJL^{+}12, WZH^{+}07, WSTI07, Wei97, WAR94b, WvdH03, WPM06, WPM07, WBD08, WMdLA10, WG11, Wii05b, Wil99, WPH03, WHR10, WPZA11, WK06, WVDS05, WCGB05, WKD08, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, YNB08, Yal10, Yao95, YKF97, YLWY06, YJKSW10, YC11, YSL12, YMY^{+}10, YY09, YAKM^{+}12, Ytr06, YLL09]. Selected [ZiTM10a, ZiTM10b, ZiTM10c, ZL97, ZCDT10, ZZZL12, ZLY^{+}07a, ZHP12, dCdF^{+}08, dBBHL10, dGEK11, dGN12, dKIM^{+}07, dMGFS11, dSOSR06, tCZ07, tHBP08, vB97, vDKS11, vES10, vKS12, vLS11, vdBK05, vdBGG10, vdHIP11, vdHP12, vdMS11, AMW07, ABK05, ADB^{+}05, AL12, AGL06, Bai05, BDGM05, BK05a, BBH^{+}05, BB05a, BR06, BK05b, Bir07, BC05, BL06c, BW91, BT06, BTF05, BS06d, CP06a, CHPW05, CR05a, aCF06, CZ05, Cha05b, CSV05, CKMS05, CKL05, DS05a, DLT05, DS05b, DPA05, DL04b, DRS05, DOSY05, DLW06, Dur12, ELM05, EW05, Eta05, FV05, FRA05a, FLS06, FRS05, FFRS05, FMO05, FPMW06, FKKB^{+}05, Fri04, FLMS05, GS05, GT87, GT90, GN05, GHMT05, GT05, GF12b, GAN05, GM05, Gue05, GRR05, HS12b, HH04, HSSY05, HPR05, HRTR05, JSdSA05, JNS05b, JY14]. selected [KK87, KIRK05, Kim05b, Koh06a, KS05a, KS05b, Kra05, KBC05, Lag05, LVW05, LBO05, LY05, LMT^{+}04, LA05, MRK05, MNS05, Mar05a, MPPJ^{+}05, MTW05, MSS05b, MZWZ05, NOOS05, OGM05, OFK05, Pac05, PTRH05, PDH99, PDH01, PDHS03, PC05, Per05, PSO05, Pie05, RMR05, ŞS05a, ST88, SVV12, SHGB05, SDH05, SLB05, SKY06, SS06d, STLK05, TLC06, TBG05, TYNM05, TAS05, Van05, VT97, WPM05, YBM87, vEHD05, AMW07, AKS09, BY07, BGS11, HH04, HA00, JRSN09, JY14, MZ04, MV12a, NH03, PT06a, ST01, TM99, VY01]. Selection [SGG06, MDBF01, MRS97, PMD90, PMD92, PMD94, PMDA96, PMD97, PMDK00, She85]. Self [BSH06, BHJY07, CdLG^{+}09, LH11, LRS03, PM09, SKRZ04, BSA05, BJM^{+}05, BG11b, BK04b, DH01, HT05, HR08, HH03, HS08c, HK07, KMF06, KH12, RSL01, SH09, WMdLA10, dMS06]. Self-Adaptive [CdLG^{+}09, LRS03, RSL01].

Self-Managed [KMF06]. Self-Managing [BK04b]. Self-Organising [BSH06, BHJY07, SKRZ04, BSA05]. Self-Organizing [LH11, PM09, BG11b, HS08c, HK07, KH12, SH09, WMdLA10, dMS06]. Self-stabilizing [DH01, HH03]. Self-star [BJM^{+}05]. Self-Sustaining [HR08]. SelfMan [KMF06]. SEM [GM05, CPvdH03]. Semanic [AGS^{+}11a]. Semantic [ACN^{+}07, AST05, ABMS06, AB04, AGS^{+}11b, AGS^{+}11a, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, AWA^{+}11a, AWA^{+}11b, ABW10, BHHK08, BCBG05, BHM^{+}04, BKH^{+}09, BHM03, BHM^{+}02, BDFS04, BFOY04, CPSA10, CKM^{+}09, CDA^{+}06, DGG^{+}11a, DCMS09, DA08, DHH^{+}08, EK12, FS05, FSA^{+}07, FSM03, FKM07, GCFA12, GMBM05, GPB02, GPE05, GPYD09, HH02, MPvH04, MSG06, MSS04, NKC09, OS04, OPS11, PST11, PCK^{+}12, PSPH^{+}10a, PSPH^{+}10b, PdA^{+}11, RDK09, Sch98b, SW03b, SSD^{+}08, SCP^{+}12, Som09, TFE^{+}09, dCdF^{+}08, Kum90, Uhl82, AKSO06, BM09b, BTF05, CS05a, CCG08, CKM^{+}09, DHRT10, DGG^{+}11a, DHH^{+}08, FSA^{+}07, FMPT10, SD06, Vis04]. Semantics [AKSO06, BDPS02, BDN08, BGKS04, CLB11, CS12b, DCMS09, FRL07, HKM^{+}07, JRL09, KHK^{+}08, KVMH09, MSG07, SMA03, Spa03, SBJ^{+}05, SAC^{+}06, SACM06, Spa06, SAF^{+}07, Spa08, SPT^{+}08, Spa09, SZS09, SD09, Spa11, AF99a, Bli87, Bro92, Bro94, Gre75, Mai88, Mai90, Mel86, PB92, Rus05, SDM05, Vog92, Zil87, ZS05a, ABG^{+}06, Ber03, BGKS04, Gue90, Jon80, Kah79, KMP84, Klo87, MR12b, Pal09, ST08, ST11, Tah00, Tah01, TL98, dBdRR93, Don76, Gue81, Pet99]. Semantics-directed [Jon80]. semantics4ws [ED06a, tHBP08]. Sematics [RCLE05]. SEMCCO [PDSD10, PSDS11a, PSDS11b, PDSN12]. SEMELS [MTH08, MHD09]. Semi [Urs02]. Semi-Automatic [Urs02]. Seminar [AFKL04, BSW95b, BMS04a, CRYS09, CMOZM11, CKLM09, DGT96, DO89a, DO89b, DMWW88, EEK^{+}87, EKLM87, Fre98, Gun94, HHLR08, IPSK04, JLM00, JKB96, KEM07, MWD06, MBS05b, PJ97, Pla96, RSS06, RHD08, RME04, SKP07, SS76, BHGP02, Die02, KSHS01, LBCO04, MM82, MB87, BRW85, SS76]. Semirings [Bis04]. Semistructured [CM00, Qui04]. Semmering [FK95a]. SEMPER [LPSW04]. SemWAT [RBsC^{+}06]. Sendai [AI97, BKV10, CK06c, CTS07, HM94, HST07, IH90, IM91, IY95, KP01a, LM97, NS81, Tok07, vLWH^{+}00]. SenDe [WZH^{+}12]. Senegal [CV10, NP11b]. Sensations [KvETvdH10b, KvETvdH10a]. Sense [KMYJ08]. Sensing [BMPM09, HLMW06, KFLS07, LKK10, RLT^{+}08]. Sensor [ASE^{+}07, ASAM07, BGW06, CSJD06, DPLW08, Dol09, ENO12, Fek08, GVO^{+}10, GAA06, JWH06, KWW04, KSHM09, KCK08, LV07a, MVCM10, MW11, NR04, NR06, NCJK08, PH12, PISW05, RMDS10, RS09b, Sch10a, SKB10, SMCM07, Ver08, WW07, ZOCJ07, CHPW05, MTW05, RKM06, ZG03, CBN99, Hag02]. Sensor-KDD [GVO^{+}10]. SENSORIA [WH11]. Sensors [BY06]. sentences [MD96]. Seoul [AYBH^{+}06, BCDN07, CCK^{+}08, Cha02a, CPC07, Chu03, Col07, CKL05, DKY06, Dat09, Gel04, GTM^{+}09b, GTM^{+}09c, HSSY05, HI10, Hu09, JP06, KCR04, Kim01b, Kim02, KSB11, Kim12, KDWW10, LBP00, LB03, LL03, LhK03, LL07, LCHS08, LC09, LBM^{+}10, LH10, LW11, LL04, MY06, MTU09, MBJ^{+}07, NR07, OB07, PC05, PCA^{+}09, PK03, RN11, SHX^{+}06, Son00, Wha03, Won01, WK06, YMHH10, Yi06, YKM06, ZSY^{+}06]. Separation [AJRB09, DJAP07, PP04, REPH06, TCYZ12, VZM^{+}10, YM01]. Sept [Sch74]. September [AI97, AKK04, AV99, ABS01, ABMM07, ARSB09, AFvZ02, ACP01, AT02a, ABK^{+}09, AFP^{+}10, AW12, APS08, ABK05, AR04, AS03, ABBC10, ADB^{+}12a, ADB^{+}12b, AD98, APO96, AL04, APPE09a, APPE09b, ABB^{+}10a, All96, ARC^{+}07, ADR07, AFM01, ABMH08, AJB09, AC09, AMGJ12, AL03b, AC11, AYBH^{+}06, Ame91b, ABD^{+}99, AFS02, AFGvE10, AFGvE11, ACH00, ABM02, AFL03, AH98, AIL12, AAB^{+}07, AKL^{+}99, AHJ07a, AHJ07b, AHJ07c, AGM03, AÖ12, AS07b, APV09, AM02, AB06, ABT04, ABW10, AD11, AJ01a, ACJ08, AHSC11, ACHH12, AGL06, Aus86, Aus93, AKTT12, Avg12, AGCP^{+}11, AE06, AS07c, AKR^{+}09, BBE01, BN09a, BBdB12, BMS^{+}08a, Bak95, BMK08, BMK09, BBGS10a, BBGS10b, BBGS10c, BST01, BdBC01, BSC12, BS09, BJ10, BFGM05, BM05a, BMUV10, BCD^{+}03, BA99, BBF^{+}06, BGLT05]. September [BPAB07a, BPAB07b, Bar08a, BBB^{+}07a, BP12a, BHH04b, BHH04c, BKLM07, BMPM09, BCKK05, BBM^{+}08, BGK12, BDPS02, BDE12, BP07, BW03b, Bat01, BMP00, BMP01, BTQ02, BTQ03, BBP04, BPW06, BL04a, BBP^{+}11a, BBP^{+}11b, BPV04, Bec05, Bec75, Bec79, BIRT05, BCRB^{+}11, BB95, BC92, BPR02, Bel03, BP11, BS03a, BCST99, BDG04, BB01a, Ben06, BP03, BHM^{+}04, BHMS07, BLKP08, BR89, BCG^{+}11, BI03, BC04a, BGPTF10, BC06, BCM00, BGD^{+}05, BGF06, BSTL07, BU94, Ber96, BDH^{+}99, Bes07, BLSS11, BKW08, BKW09a, BW10b, BB03, BMS95, BL07a, BF00, BFP04, BTPPS05, BTPPS06, BTPPS09, BTPP^{+}12, BBH01, BP01, BH77, BPT06, BMJ88, BK02, BC05, BLKW01, BL06b, BJS99, BKLP11, BKLR12, BMN05, BTCW10a, BTCW10b, BTCW12, BKS01, BBR07, BB00a, Bor93, BGM94]. September [BBBB08, BH02, BS93, Bos01, BL10b, BHJ^{+}11, Bos00b, Bos96, BR07a, BF10, BC10c, Bou11, BCRT92, BEGP04, BEG04, BJ01b, BSB^{+}99, BP09b, BEM06, BH95, BFH98, BDGK00, BDV95, Bra02, Bra96, BC11c, BvdHPP09, BRR87a, BRR87b, BH00, BKZ05, BNZ06, BZ09, BB11, BBK^{+}11, BKW06, BHN97, BMS01, BD04, BAdVR08, BHQ10a, BHQ10b, BRP09, BB02, BW09, BV94, BC04b, BM06, Bud91, Bun96, BGMST09a, BGMST09b, BW97, BCC99, BR96b, BR07c, BLVV07, Bur90, BY06, BH06d, BS06d, BCDN07, BRS09, BGH06b, But10, BGW06, CV07, CG07, CYM^{+}11, CCP96, CL96, CP98a, CIW06, CDP^{+}06, CCD^{+}09, CKS04, CPR^{+}10, CKP^{+}12, CL11, CK04a, CK04b, CK06b, CK06a, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CC05, CFB^{+}05, CE01, CT03, CHD07, CMSM^{+}10, CP12, CS11]. September [CP10a, CO94, CBF09, CGS01, CJ08a, CIT11, CFH^{+}12, CDGW10, Cea76a, Cea76b, CP11, CD00a, CMZ99, CG02, CHCM11, CH10, Cha09, CF09a, CCdS03, CD00b, CM03b, CHZ08, Chi03, CC97a, CS10, CDCJL08, CBN99, CSV05, CCG08, CSY09, CK84, CJK88, CM09b, CGP03, CP99, CFVCF^{+}12, CHD98, CCM08, CG09, CCD^{+}12a, CCD^{+}12b, CF09b, Coh10, CÓ11, CM05a, Col07, CMT01, CFR92, CM00, CS01a, CGGO03, CYBF06, CAP08, CY09, CEM^{+}06, CKC11, CF99, CCSW12, CP05, CD01, CDND11, CSRS08, Cou93b, CS96, CM11, CMB^{+}08, CL08b, CP02, CDS09, CDM04, CGB11, CRS00, CDT11, CK93, CLBP11, COB06, CM05b, CD11, CD12, CE00, COB^{+}08, DGT10a, DGT10b, DGM08a, DGM08b, DN07, DKY06, DO02, Dam99b, DLV04, DMP10a, DMP10b, DMP10c, DMP10d]. September [DMP10e, DMP10f, DSLT08, DSLT10, DSHL11, DMO98, DMS04, DJAP07, DV05, DEKR09, DG01, DAB01, DC12, De 07a, DPZ12, DY06a, De 10b, DKK08, DZ10a, Dec00, DG09a, DEG11, Del97a, Del97b, DP11, DH11, DL04a, DMLW11a, DMLW11b, DMLW11c, DBB^{+}08a, Den85, DNS09, DJ90, Der92, DH10a, DGSN11, Di 97, DZ03, DKD05, DR09, DB09, DDI10a, DDI10b, DDI10c, DS96, DD10, DVV07, DTMH10, DS08, DBHS10, DGDM^{+}11, DZGS06, DSZ10, DKU05, Dit88, DPP97, DFN10, DW10, DPT08, DJ08, DK02a, DK02b, Dol06, DCFY10, DFS08, DFM11, DFT12, DFT09, DP97, DLM99, DKP00, DLO03, DDS02, DBB^{+}04, DGB^{+}06, DBB^{+}08b, DBD^{+}10b, DBB^{+}12b, DF06, DHMM96, DFD04, DKOZ05b, DKOZ05a, DRS08, DH08, DKN02, DH07b, DLM07, DLJ12, DFS06, DKW07, DS07, EV96]. September [ERW99, EHKP05, ED06a, EBT11, EK09, EM02, EGMW11, Ehr04, EHRT08, ERRS10, EEKR12, EBSS98, ELM05, EFT01, EK12, ECB06, Enc80, EOSW07, EPVW10, EBT07, EG10, EF12, ELS09, EK07, ENO12, EK81a, Ési06, Esp01, ERMS06, EVV^{+}11, Est99, ES00b, ED06b, EKL^{+}05, FRvKS10, FS05, FT11, FKP^{+}05, FK87, FD03, FKP99, FR95b, FMS10, FIAP01, FHRR05, FS09a, FdCLN^{+}11, FMP11a, FMP11b, FMP11c, FOGE^{+}09, FZJ01, FLP12a, FHFL06, FHLP09, FHKQ12, FTAZ06, FvdHKL06, FST09, FJL97, FHY08, FDC12a, FDC12b, FBV11, Fle02, FMN99, FR99a, FSV09, FP08, FYM12, FGK^{+}11, Fra05b, FKBA12, FC93, FK95a, FMNS06, FdS12, FT04b, Fre98, FNGW08, Fri01, FLS05, FvdHJT10, FK09, FG04b, Fur05, FKL08, FLP11, FSS06a, FSS06b, FTC^{+}10, GP05, GPAS^{+}06, GP08a, GTT04, GE08]. September [GMV05, GMV06, GMV07, GVM99, GD10, GANACBR10, GANACL11, GANACBdCdV12, GMS03, GY04, GL97, GLV09, GL10a, GNR07, GHR89, Gel99, Gen09, GLW07, GKS^{+}07, GMV09a, GKMV10, GKMV11, GKMV12, GK06, GG02a, GN12b, GM89, GS09a, GH07a, Gie08, GG98, Giu98, GHK97, GZR02, GK12, GL05, GMUW12, GP09b, GP98, GRK^{+}10, GS05, GMS06, GF12a, GPPW08, GHSY06, GLC^{+}11a, GLC^{+}11b, GTVC06, GCM08, GPS03, GKS05, Gór06, GH96, GN90, GN05, GK09, GBMS11, Gra05, GT93, GHMT05, GNTF02, GS08, GPT10, GAM^{+}11, Gru98, GO06, GH93, GMV09b, GKMN10, GN98, Gru77, GC81, GM05, GVT^{+}11, GS06, GB07, GG02b, GHMV11a, GHMV11b, GHMV11c, GJTS06, GKN03, GW00, GPV88, HP02a, HOC07, HM11a, Hac75, HM11b, HM08, HT97, HCT02, Ham08, HT09, HMT10a]. September [HLSZ11a, HLSZ11b, HT11, HHMT12, HDC99, HSJ07, HTW02, HMS09, HAB02a, HHJ^{+}86, HS05a, HR96, HHM97, Har01, HRC06, HS08a, HS94, HG96, HK98a, HZK08, HS76, HPV00, HOD96, HN06, Hee94, HL04, HLW04, HR95, HP94, HS95a, HP02b, HPR05, HMM12, HHJZ12, HBE07, HG98, Hic04, HSS07, Hil06, HRR^{+}06, HL10, HS95c, HMP99, HAB02b, Hof97, Hof01, HK96, HLW08, HW06b, HNQ08, HTMC09, Hor00, HSH08, HCDL09, HY03, HZB07, HKNW07, HYMZ11, HRA02, HLCZ11, HSSV12, HWLJ08b, HWLJ08a, HJL^{+}09a, HJL^{+}09b, HMT10c, HLFS12, HK07, IKR00, IT06, INR07, IM91, IT03, IT08, IS12c, JOD98, JN10, JK00, JLV02, JSO04, JSO11, JSWhT11, JKL02, Jay99, JS97, JNR11a, JNR11b, JZ03, JKB04, JNH11a, JNH11b, JE06, JHC02, JSK10, JP09a, JNPV10a]. September [JNPV10b, JNPV10c, JWbG^{+}06, JWG^{+}06, JK04, JLW07b, Jon00, Jon01, JP05b, JP06, JP07b, JP10, JP11b, JP96, JVA04, Jos88, Jos00, Jou85a, Jou94, dCJB12, JN12, KSG^{+}09, KG12, Kah87, KMTW03, KMT00, KWA01, KWA02, KMW03, KMW04, KC05, KM08a, KKS11a, KKS11b, KKKS11, KLZ07a, KLZ07b, KR10a, Kar94, Kar77, KK11a, KLP04, KLB^{+}06, KW07a, Kei09, KJ91, KS01, KGRD10, KS03c, KHJ05d, KLP10, KS12b, KL12, KT06, KM04, KDS01, KL92, KR02b, KJB09, KKKN05, KGG^{+}11, Klu97, KZ01, KOS02, KOK04, KRP06, KHPS07, KPP08, KLLP12, KKdM^{+}07a, KKdM^{+}07b, KK09a, KHL10, KSDO06b, KSDO06a, KW07c, KHW03, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KLPP03, KC00, KS07b, KS10a, KM00, KHS07, KU12, KFM07, Kow01, KR10b, Kra03, KU06, KLN07, KO07a, KKDV02, KKD04]. September [Kra99, KMS89, KMVY06, KSH05, KS12d, KLC07, KASS07, KNK08a, KNK08b, KS96, KMN94, KWT03, KK95, KW96, KLT09, Kut98, Kut12a, Kut12b, KPW99, KCG09, LJS08, LO02, LY04, LJR^{+}10, LPT07, LM00, LS10a, LdRV94, LSW01, LSW03, LN04a, LKD08, Lau93b, LK99, LS04a, LD97, LGTB03, LGBT03, Le 94, LLP99, LYBU10, LHŚ11, Lee11, LMSS07, Leh11, LT05, Len93, Len97, LVP04, LS05, Leu03, LA94, Lev98, LFIM07, LLIH07, LJJG07, LFJI10a, LFJI10b, LJS^{+}10, LEP^{+}10, LSTZ12a, LSTZ12b, LS06a, LDTU04, LPH00, LMCH12, Lio04, LKT08, LMV98, LA05, LLO08, LSIT11, LMC12, Lop06, LHJ08a, LHJ08b, LHJ08c, LCG97, LS09, Luo04, Luo05, Luo06, Luo07, Luo08, Luo09, Luo10, Luo11, Luo12, LS10b, LIH99, LCDT08, LDC^{+}09, MJYT06, MRN07, MOP11, MOH^{+}12, MPP94, MPK04, MT09a, MDY^{+}10, MDHB11]. September [MF11a, MF11b, Mal95, Mal97b, Mal99, Mal01, Mal03, Mal05, Mal07, Mal09, Mal11, MST12, MN02a, MPS06, MT04a, MM01, Mar05a, MRŠ03, MMV04, MVC07, MSZ09, MD04, MBH^{+}09, Mar01, MNW01, MA03a, MBS05a, MSW07, MP10b, MTCOK10, MM89, MR12a, MMOS99, Mat01, MM03, MMP05, MM07, MM09b, MT97, MP03b, May01, Maz76, MB06, MH05, MC94, MLTC11, MAB^{+}02, MHA09, MF11c, MAFS08a, MAFS08b, MB09, MK03a, MD96, MS12c, Mio93, MS00b, MSG06, MR97b, MT99b, MK01, MTWD06, MR02, MR95, MvL10, Mon01, ML95, MS11a, MMC05, MBJ^{+}07, MBF08, MHW12, MS10e, MGSM12, MSS05b, MJL02, Myc95, Nad99, NL01, NWL06, NSM00, NW01, NPPDG06, NNPW12, NP11a, NP12a, ND09, NDF94, NHJ04a, NHJ04b, NHJ04c, Neh88, Ngu06, NW97]. September [NS87, Nic86, NCBT04, Nie99, NR00, NdG04, NL99, NS98, NBC^{+}06, Nor04, NSFB10, Nor97, ND04, OROD07, OSR^{+}02, eOP05, Ohl93, OS04, OAdGBP00, Oli00, OS09, OS03a, OdQ12, Oqu03, Oqu07, OD12a, OD12b, ODV08, OV09, Pac05, PT95, PAV^{+}10, PM08, PV07, PPP07, PHJ03b, PHJ03a, PGM98, PVW10, Pal01, PVV05, PGA^{+}05, PZYS12, Pat02, Pat00, PAD08b, PW92a, PMT09, PMT10, PYLW08, PMA^{+}07, Pel07, Pel11, PH11b, PA03a, PS96a, PD04, PP99, Per11a, Per11b, PCS^{+}07, PSO05, PT97, Pet01, PBGK02, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, PNPP12, PTZ03, PVP05, PMKG07, PY05b, PNB04, PS03, Pfi00, PMDK00, Pie05, PB08b, PS11a, Pit86, PPR87, Pit89, Pit91, PBS08, PLK^{+}12, PR00, PLI08, PT11, Pri91, PR98b, Pro96]. September [Pro11, PS12c, PS11b, PW07, PW08, PSS98, PP04, QLK^{+}10, QLK^{+}11, Qui98, QS00, RF01, RWSS12, RW11, RA12, RS05, RT12, RTZ08, RBGS04, RMFG06, RCT05, RRBE09, Rau04b, RLKHL12, RR98a, RBG09, Ret97, RMS^{+}05, RT95, RK12, RGW01, RMTW11, RABR03, RL11a, RH01, RV05, RV02, RBW95, RBS93, RB10, RWD09, RF12, Ros03, RBGG06, RH98, RS01, RD12b, RJL^{+}08, RSK08, RGLA09, RBB^{+}06, RKNV09, RS09d, SST00, SBV^{+}11, SS08a, SO07, ŞSW07, SKS^{+}06, SW09a, SRG04, SYMA09, Sd98, ST88, SH12a, SCKR10a, SCKR10b, SB95, Sch04a, SKMT04, Sch93b, SKP10, Sch06, SB07, SDR^{+}00, Sch10b, SJW^{+}12, SC11, SWAF95, SM06, ST12b, Sco02, SVS08, Sel95, SG10, SSJR10, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, She01, SPON08, SM11b, SMPU98, SB12a]. September [Ska01, SYP^{+}05, SBG04, Smi01, SC98, SKP00, SKP02, SKP04, SKP06, SHKP08, SHKP10, SHKP12, SPSV^{+}04, SL99b, SOP12, SK97, SDP97, SZ00, SBM11, SP93, SEN07, SEN08, SPB00, Spi95, SSU12, SHL08, SA97, SSv94, Ste94, SW97, SP07, SS09d, SCKR03, SHS10, Stu08, ŠPTM00, SZ11, SBH07, SBH09, SZT^{+}08a, SZT^{+}08b, SWSY11, SM09, SHO99, SBH^{+}04, Tah00, Tah01, TWQ^{+}12, TBE08, TMG10, TMdB11, TMS12, Tar91, Tau08, TC99, TdLdOQ08, TV94, TYZ11, TTM10, TD07, TFE^{+}09, TW12, TG00, TJ08, THX^{+}07, TBH03, TT06a, Tob80, TP09, Tou94, TBD12, TL02, Tra04a, Tra04b, TZS00, TMP04, Tro11, Tuc01, UMN^{+}08, VZ06, Van97a, Van02, VV03, VRHJ09a, VRHJ09b, VKJ03, VZM^{+}10, VKMT11, VDÉ^{+}12b, VDÉ^{+}12a, VBS^{+}10, VD12, Vog01, VSTN06, Vor92b, VAM06]. September [WF98, WT96, WRP07, Wal04a, WKP03, Wan98, WJS^{+}06, WYSS10, WLO^{+}11, WJL^{+}12, WL04, Wes01, WMdLA10, WH95, Wii05b, WM03, WSGA06, WSG07, WSF08, WSJT09, WCJS10, WWD99a, WWD99b, WL03, Win04, WNL05, Win78, WDKK07, WGD05, WPH03, WPZA11, WKS^{+}10, Wol07, WY98, WLRL08, WD00, WKD08, WDPW02, WDPW04, WDM06, WDKW08, WDKW10a, WDKW10b, WDKW12a, WDKW12b, XW11, XSA^{+}12a, XSA^{+}12b, XKGS12, Yah11, YRMD05, YJMU06, YHR^{+}09a, YHR^{+}09b, YMHH10, YBL^{+}04, YX04, YWRS11, YJKK06, YI91, YM01, YC11, YSL12, ZBW94, ZK06, ZBRL12, ŽL08, ZiTM10a, ZiTM10b, ZiTM10c, ZZ04a, ZJ04b, ZO10, ZCGH11, ZJXC03, ZLDB05, ZKZ00, Zil87, Zim92, ZSP10, ZQ98, ŻR99, dBM10a, dBM10b, dSG05, dFEN04, dMS06, dRF01, dRS01, dRLP98, dSADM07a, dSADM07b, dVGR04, dCHM12, tCZ07, tHBP08, vB97, vE91]. September [vG02, vKS12, vLS91, vL94a, vLS11, vdPW10, vdBK05, vdHHsHD06, vdHXMW08, vdHIP11, vdABCC05, vdMBD08, vdV84, JS05]. September19 [CPSA10, AJ01b]. September19-21 [AJ01b]. Sequence [SG01]. Sequences [GG98, CP10a, GGHS07, GPPW08, GHSY06, HSSY05, HJ12]. Sequential [MO06, AJS05, Bur97, Wal82]. sequential/parallel [Wal82]. SERA [DLW06, RLL04]. Serbia [CIT11, LPR12, Ong11]. Serena [HP02a]. SERENE [Avg12, Tro11]. Sergot [ACÇ^{+}12]. Series [DFGN12, Cha85, Lau93a]. Serious [MOP11, MOH^{+}12]. server [Sch93a]. Service [ALS^{+}11, CW11, DKK08, DY10, Hei12, Kou11, MCA08, MPP08, MRK05, MNS05, MRvM07, NMPM08, PPPM10, PRT06, Rei03, SPSV^{+}04, Sto04, WH11, dMB05, ABLM05, BKB03, Cea76a, Cea76b, Del00, KMN94, LPH00, MB01, Mar03, SS99, TAS05, WHS01, Zui99, BCS09, BCT06, BKM08a, DL06b, DGT10c, DPZ12, DR09, DLdS04, FL09, GRB^{+}07, HKM^{+}07, JSV11, KMMN11, KHK^{+}08, KVMH09, KLN07, LLTY12, MWYF10, MRK05, MNS05, MRvM07, MRY^{+}11, NMPM08, OWPY03, PJC^{+}12, PPPM10, PRT06, JSW03]. Service-Based [ALS^{+}11, CW11, DY10, MPP08, PPPM10]. service-IWQoS [JSW03]. service-orientation [TAS05]. Service-Oriented [WH11, BCS09, BKM08a, DL06b, DGT10c, DPZ12, DR09, FL09, GRB^{+}07, HKM^{+}07, JSV11, KMMN11, KHK^{+}08, KVMH09, KLN07, LLTY12, MWYF10, MRY^{+}11, OWPY03, PJC^{+}12, BCT06]. Services [AK10, AGB^{+}11, ASP01, AH02, AR10, AH07, BG09, BS03a, BPZ09, BPS06, BAP^{+}05, BKZ05, BNZ06, BB11, BKM06, BKM08b, BW09, BCF^{+}02, BHM^{+}02, BFOY04, CL11, CP12, CS05a, CGS01, Cha02a, Chi03, CCBW11, CGS07, CRS00, Del00, DCMS09, DH08, GCC^{+}02b, HS08b, HJMP06, HTMC09, Jac07a, JZ03, KR10a, KS03b, KMF06, KT06, KPR07, KMN94, LS10a, LSgL^{+}06, MA03a, MP10b, MKR^{+}04, PAD08b, PNPP12, PB09, PvS07, RS00, SP09, SNČ^{+}12a, SvS00, Sch02, SSN08, Smi03, SPSV^{+}04, Ste07c, Ste09c, Ste11c, SSKR02, SD10, VAR12, VFC08, ZJ04b, dMS06, AMU03, BS06d, BMP96, Cha03a, CCK95, DH05, DD97, DOS99, FM97, GP98, Hal96, HS98d, JKB04, Kah03, KMTU03, LG99, MBC^{+}97, Pla96, SDH05, Smi01, SW97, Tri98, VH04, Zui99, ACK00, MSRP98]. Services [Sch02]. ServiceWave [ALS^{+}11, BCS09, CW11, DGT10c, DY10, MPP08]. Sesimbra [AW90, VA95]. session [Li96]. Sessions [IM10, IJ06, LS12a]. SET [HKM^{+}11, MNW08, SSZ12, YGL^{+}10, CKPB07, PFCS05]. SETA [CP10a, GPPW08, GHSY06, HSSY05, HJ12]. SETN [APS06, DVVA08, KPK^{+}10, MPV12, VS02, VP04]. SETOP [GANACBR10, GANACL11, GANACBdCdV12]. Sets [APSZ02, ASR^{+}07, CGBZ08, GHH^{+}06a, KPRS07, KŚHM11, LNW^{+}12, MMP07, PSGB^{+}04, PS05b, PS06, PS08b, PSR08, PSW^{+}09, PS10, PSS^{+}10b, PSC^{+}11, PSS^{+}11, PS12b, SCH^{+}09a, SWS^{+}05, SYP^{+}05, SKR^{+}10, TSKGB04, WPSY06, WLGB^{+}08, WLP^{+}09, YLW^{+}07, YRWS11, YYS^{+}12, BdBK03, PSD^{+}05, PS05a, PSD^{+}07, PSM^{+}07, PS98b, Wan03, ZSO99, ZY01, PSD^{+}05]. SEUS [BGR08, LN09, MPPU11, ONPR07]. Seventeen [Wie06]. seventh [Nor87, vLdB80]. Sevilla [CDP^{+}05]. Seville [GRT02, HAB02b, MMG10, OSC^{+}08, RDPMA^{+}11a, RDPMA^{+}11b]. SFM [ABDW10, BB05b, BH07, BDZ08, BPZ09, BI11, BCP12, CB06, BI03]. SFM-MOBY [BB05b]. SG [BFKO05, BFKO06, BFK^{+}07, BFK^{+}08, BFC^{+}09, DVM^{+}11, BKO03, BKO04]. SGC [THX^{+}07]. SGCA [CPGL05]. SGDA [MOP11, MOH^{+}12]. SGS [CAS^{+}09b]. SGT [JPXS04b]. Shafi [Gol11b]. Shai [Gol11b]. Shan [CWC^{+}07, DLW^{+}07]. Shanghai [ACL^{+}04, BBCL08, BRL08, yCCZ07, CLWC08, CCM^{+}04, DWWL09, DZ10b, FX08, GM03, GLW07, GPYD09, HSSV12, HWLJ08b, HWLJ08a, JE06, JL97, LC06, LYL11, LSDN04a, LSDN04b, LOS05, LFIM07, LLIH07, LLWL09, LDXZ10a, LDXZ10b, LLTY12, LZ00, LSU07, LZK11a, LZK11b, LZK11c, Luo07, LSW^{+}10, LCY^{+}11, MRN07, PZ08, PSPH^{+}10a, PSPH^{+}10b, Qin10, QLXL07, QLK^{+}10, QLK^{+}11, SZ11, TYZ11, Ued10, WTZ^{+}04, WPRW06, YJS^{+}06, ZHF04, ZTY^{+}05, ZLK10a, ZLK10b, ZCDT10]. Shannon [Som09]. shape [Ley01, van90, FMdGC99, Cip95]. Shared [CCJ07, MCdS^{+}08, EV01, Vos03, Cha05b]. Sharing [KZ10a, NT06, FNPW05, SDM05, GJS08]. Shatin [LwCM00]. Shefayim [EKM06]. Sheffield [BMH05, ENP02, Fog95, GAM^{+}11, Jen92, KSG^{+}09, MSW07, PPH07, WNL05]. Shelf [Mor06]. Shenyang [IHX^{+}08, WYP12a, WYP12b, ZYBX08, ZHLW12]. Shenzhen [CHB^{+}12, HCS11a, HCS11b, TSJ12a, TSJ12b]. Shibpur [RCC12]. shift [Ron84]. Shimon [GRS06]. Shiraz [ST02]. Shizuoka [Kur10, LMMN10]. shooting [dB93]. Shores [KKP02]. shortest [Akm87]. shrink [JRR03]. sic [AS02b, CJ95, GGHN97, MP93, Pau93, STT04]. Sicily [Cra12, MTS^{+}03, MT^{+}03, NCBT04]. SIDE [RKS^{+}12c, SH12b]. Side-Channel [SH12b]. siècle [GT93]. Siedlce [KY07b]. Siegen [PRdRW09]. Siekmann [HS05b]. Siemens [HRR93]. Siena [BCKK05, CLM03, CLS07b, CLP05, GLN01, vdL04, vzGIK08]. SIGIR [CCS04]. SIGKDD [BFMS08, BFJM09]. SIGLEX [PB92]. Sigmetrics [DN93]. Sign [EKF12, WF98, WS02]. Signal [AJRB09, DYCD09, DJAP07, ELNM08, ELN^{+}10, EML^{+}12, PP04, REPH06, SPO11, TCYZ12, VZM^{+}10, WR09, Per06, PVC^{+}99]. Signal-Image [DYCD09]. Signals [EHMM09, PS07a, PS08a, ÜÇA10]. signature [Pfi96]. signatures [Pfi96]. Significant [dRLP98]. SIGPLAN [BCT02b, DM01a, MB98]. SIGPLAN/SIGSOFT [BCT02b]. SIGSOFT [NL99, BCT02b, JS97]. Sigüenza [GPB02]. SIIS [BKL^{+}12c]. SIL [Hou90]. SIM [YYH^{+}12]. SIM3 [XYZU11]. SIMBAD [PH11b]. Similarity [Yu02, BHVV09, NP12b, PH11b]. Similarity-Based [BHVV09, PH11b]. SIMPAR [ABH^{+}10, CNP^{+}08, NABK12]. simple [MP95b, Eng74]. Simulated [McK99, BOH^{+}12, DBC^{+}10, LKZ^{+}08, McK99, WLC^{+}06, YKF97]. Simulation [AT07, APN08, AHMT08, AGCP^{+}11, AS07c, BE08, BC10a, BDW08, BDSW09, BGJ11, CPR^{+}10, CM04b, CS12c, DS09a, DP10a, DGWB09, DGG^{+}10, GCK06, HENR03, HPTV05, HS06, HAB02b, Kim05b, LFIM07, LLIH07, LFJI10a, LFJI10b, LJS^{+}10, MPK04, MvL10, NBC^{+}06, SA06, SPB00, Squ09, SM09, VWH06, VBH07, VSMS12, VAM06, WPRW06, ZBH12, vLS11, AD03, Bai05, CM03b, DM01b, DLT05, DTV02, Hou90, Lau93a, PVV05, PV04, SCG98, SBD03, ABH^{+}10, CNP^{+}08, LWZ09, NABK12]. Simulations [BBDS12, DBSvD09, Dig11]. Sinaia [PLK^{+}12]. Singapore [Abe10, BDZ04, BLO^{+}07, BSZ12, BC10c, BHH^{+}06, BY06, CCD^{+}06a, CCD^{+}06b, CWY^{+}06, CEK11, DPLW08, DSZ^{+}11, Dec00, DQBZ02, DBPZ05, DW03b, GLCT10, GGK^{+}05a, GGK^{+}05b, GGK^{+}05c, GGK^{+}05d, HMCL06, IZ02, Ish99b, JY96, JP10, Kam99, Kap08, LOX99, LM98b, LTW06, LYBU10, LIJ97, LLC^{+}05, Li96, LYT06, LSS^{+}04, LFK^{+}02, LXH^{+}12, LMV95, LRL98, MSWZ06, MSH05, NP11a, NZ06, NKLC06, NLKJ06, OA93, RSV07, Sha07, SJT05, SXC05, VW08, XXX^{+}09, ZLDB05, ZYB06]. singularities [OFK05]. Sino [ZZZL12]. Sino-foreign-interchange [ZZZL12]. SINOBIOMETRICS [LLT^{+}04]. SIPEW [KGS08]. Sirince [PSE10]. SIROCCO [EH12, FG06, KY11, KŽ10b, PSE10, PR05, PZ07b, SF08b, KS04c]. SISAP [NP12b]. Sitges [Com97, GVCA07, MCP93, MMC05, SB95, vB05]. SITIS [DYCD09]. Situation [KM11]. Sixteenth [MSY00]. Sixth [MP01a, MP01b, BF89, Mau79, Nor86].

Skills [JNS05a, RLKHL12]. SKLOIS

[FLY06, LYL06, PYLW08]. Skövde [GB97]. Sky [WZ03]. Skye [GL03]. SLE [GLV09, MBS11, SA12, vdBGG10]. Slovak [HS98c, Jay99, PR01]. Slovakia [ABH12, ČGH^{+}11, GKB^{+}08, HPZ10, KM11, KU06, KN09b, MN02a, NR00, PRR94, PR97, Rov98, RV03, RSW12a, VBCBS05]. Slovenia [BSTL07, BG12a, BGMST09a, BGMST09b, D’H10b, DLŠ11a, DLŠ11b, DF99, ERW99, EF12, GKMV12, KŽ10b, PLC11, RT12, RtTM12, SGG06, SL99b, TW12, dGN12]. Slowakia [KS04c]. SLS [SBH07, SBH09]. SLSPS [SGG06]. Smalltalk [De 07a]. SMART [CW11, AGB^{+}11, ABK12, AJ01b, AN06, BMK08, BMK09, BDK10a, BKH11, BGH06a, CCNY10, CS12c, DPNM12, KR10a, LBM^{+}10, MBH^{+}09, MKB^{+}09, OYM07, OSC^{+}08, PPH07, STP^{+}10, SMBQ07, SKM07, TBKO10, AJ01a, BBH^{+}05, HG96, AJ01b, BMK09, BDK10a, BKH11, BMPM09, BKO03, BKO04, BFKO05, BFKO06, BFK^{+}07, BFK^{+}08, BFC^{+}09, DVM^{+}11, DFPS06, GLIC10, GS08, HLMW06, HMW^{+}08, KFLS07, LKK10, Pro11, QS00, RLT^{+}08, TBKO10]. SMILE [Pol01, KvG98]. Smokovec [ČGH^{+}11, GKB^{+}08, KN09b]. Smolenice [DK90, HS98c, KJ91, KS04c]. SMVP [CKMS05]. SNA [WJL^{+}12, ZSM^{+}09]. SNA-KDD [ZSM^{+}09]. SNAKDD [GSYZ10]. SNePS [Kum90]. Snowbird [GQ11, GM11b]. SNPs [IWC04]. SNSC [WL03]. SNSM [YYH^{+}12]. SNSMW [XYZU11, YMY^{+}10]. SOAR [WMdLA10]. SOC [MRY^{+}11, PJC^{+}12]. SOC-LOG [MRY^{+}11]. SOCASE [HKM^{+}07, KHK^{+}08, KVMH09]. Soccer [AK99, BLNG10, BCT02c, BJNT06, IMWZ09, KLR03, LSST07, NRSSV05, PBBY04, RMSS12, dSCP11, SBK01, VPK00, VROD08, Kit98]. Social [AHSC11, ACHH12, BHMS07, BBK^{+}11, BSPB07, CSM10, FFM05, GLCT10, GSYZ10, LMP04, NKC09, OZ09, OZ11, PS09c, Sch07, Squ09, ZSM^{+}09, vEV12, vdBK05, CW94a, Chu03, HWP01, Ish99a, Oss99, SILS05, BMW10, DSZ^{+}11, GLCT10, GKC^{+}12, MBH^{+}11, SYNC11, YGE12]. Societies [ADP09, AOSV08, APV09, CKLN10, DGR06, FBSMS08, GOZ05, HNS^{+}12, OROD07, OPT01, OPP04, PTZ03, SSO^{+}11, FST01, FBSMS05, OTZ01, Oss99, Pad99]. Society [Ber08, CS02a, CS12c, Ham00, Hol07, Hol08, HM09, HS11, KL05, Kim01b, KW07b, LM06a, LS09, LCDT08, LDC^{+}09, McC96, MN98, RS00, SS04, SM01, TG04, VU09, XCd03]. SocInfo [BMW10, DSZ^{+}11]. socio [Kim01b]. socio-economic [Kim01b]. sociological [TvdBI02]. Socionics [FFM05]. Sociotechnical [ZPT^{+}06]. SOCO [YCC^{+}08]. Sofia [CKD11, LP10b, LNSS11, OF15]. SOFSEM [BSW95b, BFG^{+}12, ČGH^{+}11, GKB^{+}08, GP02, HJW00, NKM^{+}09, PR01, Rov98, VPBŠ03, VBCBS05, WTP^{+}06, vLIvdH^{+}07, vLMP^{+}10]. SOFSEM’96 [JKB96]. SOFSEM’97 [PJ97]. SOFSEM’99 [PTB99]. Soft [BS11e, Bis04, MCA^{+}07, ORB^{+}09, RSTZ04, RTZŻ06, RTZZ08, RST^{+}10b, RST^{+}10a, RKS^{+}12a, RKS^{+}12b, SAG10b, AD03, AAN00, NA97, PS02a, ZSO99, BLS02]. Soft-ware [BLS02]. SoftSpez [EDD04]. Software [AK07, AL11a, AMS06, AM91b, AM91a, AG09, AI06, AS02a, AM03, ABDW10, AFKL04, AL03a, Ama08a, AT12, AS07a, AS10, AÖ12, AS12, AKG06, AP07, ABW10, BVO10, BL12, BH06b, BHKR11, Bar11, BLS12, BRS^{+}07, BPR08, BHS07, BM11a, BDG12a, BI03, BH07, BDZ08, BPZ09, BI11, BCP12, BFGL10, BF11, BR97, BZU07, Bir12, Bod05, BI04, BKS05, BCSW10, BC08, BE07, BBPP09, BC09b, BBK^{+}11, BP12b, BKM06, BKM08b, BHLP07, COBV11, CJ11, Can12, Cas09a, CSW06, CST00, CPV03, Cer05, CSR08, Cha09, CW09, CH09, CS10, CB06, Cle99, CÓ11, CDSS07, CW03, CSSW04, DR02a, DSB04, De 07b, DG02, Deg03, DG12, DO89a, DO89b, DJJ12, DS11b, DP12, DW03b, DZ10b, Dro08, Due04, Dun09, DL07, EB04, ELS12, EK09, Ehr85a, EEK^{+}87]. Software [EKLM87, EDD04, EWZ11, EM10, FLVS11, FI08, FK12a, FP08, FdS12, GH03, GM89, Gho10, GO11, Gie08, GH05, GMO03, GGS11, GH09, Gor10, GHC^{+}06, GS00b, GM04b, GW06b, GM11b, GH07b, GRP10, GS06, GB07, Gup10, HZ05, Han98, HMV01, HNRW06, HMP08, Hed03, HKP10, Hen08, HM03, HP06, Hof11, HCR06, HI10, Hor02, IT06, IJ06, Jäh99, JS97, JS08, JP04a, Joh03, JV06, JP11a, JMP12, Jou11, KP00, KS02c, KOFJ09, KMK10, KMSJ11, KP04, Kie12, KR02b, Kno11, KMS06, KV08, KK09b, Kos98, KP09, KO07b, Küh07, LSV06, LVS08, Le 02, LOR10, LCH^{+}12, LKK^{+}07, LPH09, Lin01, LH06, LMA08, LS04b, LM04, LP08, LGS09, MY01, MRZ09, MWZ10, Mei08, MH04, MR08, MNW08, MW08, MN12a, MN12b, MGH09, Mor06, MZ06a]. Software [Mül10, MV06, MA07c, MYS10, MZ06b, NZZ11, NS87, NL99, NBM09, Nyb08, O’B12, OO12, OKS02, OS07, Ong10, OSRS07, PZ07a, PR08, PPCS12, PR03a, PP12a, Păs09, PU04, PVTG07, PSM10, PH06, PT91a, PT91b, QQ11, RR08, RRR99, RS06, RA06b, RA07a, RA07b, RO09, RRMP08, RV10, RD12a, Rob06, RV11, RT10, RM04b, Sag05, San01, SVS02, Sas05, SPH07, SB95, Sch11, Sch04b, SCHS07, Sei07, Sei12, SKRZ04, Ses06, SW08, Smo00, SP93, Ste98, SHUH08, Swi99, SSZ12, UBW06, Val06, VW05, Wal84, Wal04b, WPRW06, WPR07, WPR08, WGMP09, Wat00, WMS04, WG11, WP10, Wil01a, WFB12, WKB04, WH11, WZ11, WCGB05, Yor08, dA09, dLZ12, dMS09, vEV12, vF91, vdPW10, AI97, AN95, AFM01, And94, AR98, Asp98, ABGP05, BR03]. software [BL94, BMH05, BC04a, BB05b, BJ86a, BD97b, Bih97, Bir07, BN91b, Bjo80, BHL90, BS02b, BD93, BO00b, BKS01, Bos01, BL02a, BJ95, BR98, BJ01d, CV96, CW01, Cow77, CS01b, CE00, DA01, Dwy01, Ehr85b, EW03, Est98, Fin99, Fra00, FP05, FvdHJT10, FM96, GJ93, GM03, GOW03, God05a, Gol96, GdlP99, Gor03, Got83, GAY86, GN05, Gue05, Hae99, HM94, HB97, HPV00, HCS^{+}05, HR99a, HK01, HM01, Hus97, Hus01, IM91, Joh97, JS84, Knu99a, KN83, KP01a, KMO^{+}05, Kun84, LRS03, Lam96, LB05, LBO05, Mad89, Mah85, Mai00, MS03a, Mat02, MM82, MNS95, Nag96, Nes95, NE02, NY93, Niv98, Nor86, Nor87, NK88, Nor90, OGM05, OZ01, PM96, Pez03, PK00, Pre95, RS97a, RA06a, RMS04, RMS^{+}05, RSL01, RBS93, RS03]. software [RB00, Rus00, Sch01a, Shy92, Shy93, Str96, Thi94, Thi95, Tho99, Tiu00, Vau98, WK02, WN96, WWC02, YZZ^{+}05, vdL98, ABMM07, ABD^{+}09, Amb01, AJ11, Avg12, BG10a, BPS11, BW10a, Bau75, BF09, Bit88, BK04a, BL10b, BGR08, Cha02b, CdLG^{+}09, CGLR05, CGG^{+}07, CPvdH03, Con97, Con00, CGB11, CGBDA08, DF09, Dei90, Der92, DKW99, DH94, Die02, Din04, DLW06, DBB^{+}08c, EHH12, EW91, Est95, FMS02, For88, Fra00, FMY04, Gac02, GLZ^{+}03, GCL^{+}06, GLV09, Gib89, Gru98, GO06, GAN05, GM05, GDGB12, Hei12, HKM^{+}11, Ibr95, IJ06, JS97, KC02, Kra03, LN09, Lon90, LS06b, LGR^{+}04, LV07b, MBS11, MJ07, MPPU11, Mon96, MO05b, MBF08, NL99, Nor04, NA97, eOP05, ONPR07, Oka03, Oqu03, OWM04, Oqu07, OSRS07, Pas02, PT08a, RLL04]. Software [RAM05, RRM06, SB95, Sch95, Sch04a, Sle92, SA12, SP93, Som96, TC95, Tom91, Tro11, WPRW06, WPR07, War94a, WvdH03, WBHR08, vdBGG10, vdL01, vdL02, vdL04]. Software-Architectures [BPR08]. Software-Intensive [WBHR08]. solution [ZWS81]. Solutions [DPNM12, LY01, PPPM10, Rei99, Rei94, JN01, NRW04, ALM95]. Solving [AFR^{+}04, ABFR07, Bis04, BJ01c, CCD^{+}12a, CCD^{+}12b, FRS08, FPFR05, HCF06, LO11, MA03c, MÁ07b, O’S03b, OFR09, RJL^{+}08, RBB^{+}06, SCKR10a, SCKR10b, CSV05, DSM94, EV96, EBSS98, LPA00, MAB^{+}02, Oss99, SDR^{+}00, SM91, YBL^{+}04, Bil97, Fen00, FP96, Fer98, Han02, ZZ04b]. some [ON80]. Sonthofen [BGV01]. Sophia [BF08, Ber00, BJN03, CSRS08, CDT11, DL06a, Dit01, FLS06, FZJ01, HW03, HSMPA09, HDLM09, JVA04, KMP84, MMM01, SRG04, dBBM09]. Sophia-Antipolis [BF08, CSRS08, HSMPA09, KMP84]. sophisticated [San80b]. SOQUA [RMS^{+}05]. Sorrento [CC05, DKD05, Fic03, GYD^{+}06, MM82, M^{+}06, TC95, YRMD05]. sorted [SS89]. Sorts [BHR90]. Sound [MRC12, YKMJ09, YAKM^{+}12]. Sounds [KMYJ08]. Soup [FGH98]. source [Yel88]. source-to-source [Yel88]. Sources [Urs02]. South [ABGL09, BL10a, BL10b, BBK^{+}11, CP11, FKLY08, Ho12a, Ho12b, HTMC09, HJL^{+}09a, HJL^{+}09b, LW11, gLPZ^{+}12a, gLPZ^{+}12b, LW12, MS11a, OB07, Sch11, WW03, YJKSW10, YYH^{+}12]. Southampton [Bla03b, GHK97, HHM97, PR98b, QCDM06, SG07]. Soviet [Coh92]. Sozopol [Giu98, LMWY04, LMW06, LMW08, LMW10, LMW12, MWY01]. SP [MS00c]. Space [BtHRBB12, FNGW08, Ker01, SMP07, TMLL09, ter97, GL03, KSW05, Ahn04, FCF92, Ker01, Lis89, Nie99, Sze94, ter97]. space-time [Lis89]. Space’99 [Nie99]. Spaces [ABK12, BMK08, BMK09, BDK10a, BKH11]. Spain [ABB^{+}11, ABFH09, AAB^{+}06, AV08, ANF08, AS06, AFGvE10, AK12, AMM99, AHSC11, ADL09, AGCP^{+}11, BYdVZ^{+}12, BLS06, BBGS10a, BBGS10b, BBGS10c, Bar08a, BHJKPG06, BMSMS04, BH10, BGP11, BW10a, BBP04, BHHK08, BR97, BvDMST11, BBRS05, BBC^{+}08, BCC07, BK04a, BR07a, BAH09, BHV11, BHQ10a, BHQ10b, BL05c, BFC^{+}09, CPS05, CSPC09, CRJ11a, CRJ11b, CDP^{+}05, CRL01, CS07, CO94, CP10b, CGMP11, CA08, CGP02, CAS^{+}09b, Cho12, CH99, CKM93, Com97, CUdlC04, CNmWH^{+}05, CYBF06, CAP08, CWO^{+}09, CY09, CRS10, CSA^{+}12a, CSA^{+}12b, CEKR02, CvH07, Cra08, CMM01, CRW^{+}12, CGBDA08, COB06, Cue03, DH05, DF01b, DDHP05, DAB01, DIMV09, DIMV11, DMV12, DGM09, Des09, DGSN11, DAC^{+}12, DR81, Dia83, DO89a, DO89b, DS96, DPA05, DJRZ06, DPA07, DF96, DJJ12, DGDM^{+}11]. Spain [DZGS06, DFT04, DFPS06, DPMP^{+}11a, DPMP^{+}11b, DLM99, Dor02, Due04, EOE^{+}07, EO94, EP06, EWZ11, EV07, ETG02, FKP^{+}05, FZJE^{+}05, FZJE^{+}06, FSdlPT11a, FSdlPT11b, Fer08, FIAP01, FMO05, FOGE^{+}09, FHW12, FR99a, FP05, FRSH05, FG10, FN12a, FG04b, GG05a, GGD09, GTT04, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GVCA07, GMS03, GB99, GRT02, GVB12, GN12c, God05b, GPB02, GdlP99, GTVC06, GM04b, GHP^{+}08, HMT10a, Han09, HM12a, Her00, HT05, HP94, HP02b, HM02, HC11, HAB02b, Hof03b, HVT12, HP89, HdMKS12, IT06, JL08, JP04a, KN09a, KMW04, KMD02, KLP04, KLS10, Kei09, KLP10, KOS02, KCMO11, LSW03, LO11, LTÅ^{+}11, LG99, Leu03, LA94, LV06, LSYO08, LK04, LS04b, LQO04, LSF07, LH08, LT11b, LP00, LFL05, LGM11]. Spain [LRS98, Luo04, Luo05, Luo06, Luo08, Luo10, LMB08, MRC12, MMD05, Mag02, MPP08, MMR07, MOBS06, Mar11c, MRU12, MVMP^{+}04, MBMS07a, MBMS07b, MOFM10, MVOF08, MBVTY11, Mau96, MSS03a, MAB^{+}02, MMG10, MCP93, MS95, MMMD97, MdPA98, MSA99a, MSA99b, MP01a, MP01b, MA03b, MA03c, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, MOT12, MAA91, MMM05, MSK^{+}12, Mor95, MDP04, MDBF01, MDPQA09, MDPQA12a, MDPQA12b, MMC05, MSX10, MOTN07, MMM11, MGI^{+}11a, MGI^{+}11b, MGI^{+}11c, MGI^{+}11d, MGI^{+}11e, MC05, NDB94, NL01, Nao07, Nie03, NMP^{+}04, OAdGBP00, OP97, OS09, ORB^{+}09, OSC^{+}08, Oqu07, PMT10, PA03a, PvES12, PP12b, PH02, PCPS03, PD04, PF06, PF08, PF10, PFM12, PCG^{+}07, PTZ03, PMD90, PMD94, PMD97, PB97, PLMV12, Pri91, PSS98, PP04, RV10, RDPMA^{+}11a, RDPMA^{+}11b, RW06, RBTD04]. Spain [RCS10, RGLA09, SPCG07, SB95, SSRA98, Sch04b, SS05b, SS06a, SG10, SSJR10, Ser12, SCD^{+}10, SCZ10, SPSV^{+}04, SPC^{+}08, SQL10, Sur02, TJ08, TLJ06, TN04, TNVDF06, TN08, TG10, Tra04b, TGAH11, Val94, VY93, VDPD96, VMBMN04, VBS^{+}10, Vit95, Vit10, VSH11, WR01, Wal04b, WMS04, WTM^{+}02, WKS^{+}10, YAKMJ11, Zui99, dBBC02, dFEN04, dRWK^{+}10, dPMA98, vB05, vdHS10, vdL98, vdL01, vdL02]. Spanish [BCC07, CUdlC04, FN12a, LGM11, MOBS06, MRU12, MMG10]. sparse [Col84, OZ83, ZWS81]. Spatial [BKLM07, FHW98, FNGW08, Had03, HSB^{+}10, MSP^{+}09, MEB05, PZK07, PTM^{+}11, Ren02, SSU12, AS99, Buc90, EH95, Fre03, FKKB^{+}05, GS91, GPL99, Her94, Jen01, LCS91, OA93, RH01, SV97, BKLM07, CM05a, EGMW11, FC93, FK95a, FHW98, FM99, FBHW00, Fre03, FKKB^{+}05, FNGW08, HF97, HSB^{+}10, HCDL09, KWT03, Mac06, Mon01, Sch97b, SSU12, WDKK07, dPS95]. Spatio [Kam99, FCF92, RH01, BJS99, DFGN12, Jah93, KSF^{+}03]. Spatio-Temporal [Kam99, FCF92, RH01, KSF^{+}03, BJS99, DFGN12, Jah93]. SPC [CPPB06, HU05]. SPDP [FR98, Rol98, Rol99]. SPDP’99 [RF99b]. Speaker [Mül07a, Mül07b]. SPEC [KS09a, KGS08]. Special [BS10, GTM09a, GTWC09, GTA10, GTM10a, GTM10b, GTSS11, GTM11, GTP12, HKW11a, HKW^{+}11b, HKW^{+}12b, JvdA09, KOFJ09, MKZ02, MKZB04, MKZK06, MKZK08, MKZK10a, MKZK10b, MKPZ12a, MKPZ12b, NJA^{+}11, PBG^{+}10, PPdV12, SK12, Bos96, GHK97, MO92a, ZVU99]. Specific [Tah09, LBCO04]. Specification [BM03b, BDGK00, DB07, DHM99, EDD04, FLUV02, GH06, GP08b, JJF03, JK06, Mos04, Mül02, Pal09, ASP85, BBP89, Ber02b, BBKW03, BJ86a, BKL^{+}91, BC93, BH95, BHT97, BFH98, Bre91, BMS96, Ehr91, EO94, HOD96, HKB93, Joh01, LB90, PP01, Pol81, RTA95, ST88, She85, Sta82, TKHS05, YY01, Hoo91, KN83]. Specifications [EKLM87, Bjo80]. Specify [CDE^{+}07]. specifying [ABL96, Koy92]. Spectrum [BBB^{+}85, CN06, LBO05]. Speech [Amt99, ADM^{+}08, CHG^{+}07, CEFZM05, CBS02, EV09, FZJE^{+}05, FZJE^{+}06, HM11a, MM03, MM07, MM09b, PCG^{+}05b, RG03, SRS03, SKP00, SKP02, SKP04, SKP06, SHKP08, SHKP10, SHKP12, SCZ10, SFZE07, TGAH11, ÜÇA10, CW94b, Gor02, Kom97, MMOS99, Mat01, MMP05, YAKM^{+}12]. Speech-Based [ADM^{+}08]. Speech-Triggered [RG03]. Speed [CZ02, ML04, Ham89, Ste94, FLL03]. spelling [Pet80]. Spikes [MSY08]. SPIN [BE07, DP12, God05a, GM11b, HMP08, Păs09, Val06, vdPW10, BR03, BL02a, Dam99b, Dwy01, GM04b, HPV00]. Spindleruv [BFG^{+}12, NKM^{+}09, vLMP^{+}10]. SPIRE [ATM09, AM04, CBGCCZ12, CL10, CN05, CFS06, GSS11, KTH09, ZBY07, LO02, NdMO03]. SPLC [BL10b, eOP05, Cha02b, Nor04]. Spoken [HMCL06, LMMN10, MML97, LMMN10]. sponsored [DO02, GW96, WF77]. Spontaneous [GANACBR10, GANACL11, GANACBdCdV12]. Sports [GMUW12]. SPR [dVLKR^{+}08, YKF^{+}06, CAD^{+}02, FIAP01, FCD^{+}04]. SPR’98 [Ami98]. Sprachen [BI73]. Spring [BGHJ07, Cho88, CJ95, CCR86, Gib12, GP89, Gue90, JM95, JMO10, Pin89, Pod95, BJ86a]. Springs [CS05b, DG05b, HMRT05, Kap92, Mur99, Wil99]. SPW [LBO05, WPRW06]. SPW/ProSim [WPRW06]. SQL [Has96]. Squares [Ahn04]. Sri [MVCM10]. SSBSE [CÓ11, FdS12]. SSC [GGHS07]. SSD [GS91, Buc90, EH95, OA93, SV97]. SSD’99 [GPL99]. SSDBM [AB12, CFB11, GL10b, LM08b, Win09, Mic90, RKS89]. SSDGP [BGHJ07]. SSDSN [THX^{+}07]. SSGIP [Gib12]. SSPR [dVLKR^{+}08, YKF^{+}06, CAD^{+}02, FCD^{+}04, PpWR96]. SSPR&SPR [GHI^{+}12, HWW^{+}10]. SSPR’2000 [FIAP01]. SSPR’98 [Ami98]. SSR [EK81a]. SSS [DG06b, DPV11, DCFY10, GP09c, HT05, KS08b, MT07a, RS12, HH03]. SSTD [MSP^{+}09, MEB05, PZK07, PTM^{+}11, Had03, Jen01]. SSVM [BtHRBB12, SMP07, TMLL09]. SSWS [MTH07a, MTH07b]. St [ABK12, BW05a, BMK08, BMK09, BDK10a, BKH11, BFF^{+}10, BHP95, BC92, BGU94, BKLS11, BL11, CHD98, Coh10, DDS12, Dan12, Dyc94, Dyc00, EM08, GANACBR10, GSP01, GPS03, GLS05, GKS05, GZSC07, GHH06b, HCS^{+}05, HAB98, IJ06, KS10a, KS12c, KHI06, KHS07, KV11, MWD06, Mal95, Mal99, MR97a, NM94, NH03, OS04, OWM04, PM09, RVY05, SAB^{+}03a, SAB^{+}03b, SAB^{+}03c, SAB^{+}03d, SRT06, Vau06, VZM^{+}10, Vor92a, Vor92b, Vor93, WZA10, YDW08]. stability [DXS91, Dob81]. Stabilization [DG06b, DPV11, DCFY10, GP09c, KS08b, MT07a, RS12]. stabilizing [HT05, DH01, HH03]. STACOM [CPR^{+}10, CKP^{+}12]. STACS [AF02, AH03, BVNW87, CL90, CJ91, CW88, DH04, DD05, DT06b, EFW93, EMW94, FR01b, FJ92, FM84, MP95a, Meh85, MT99a, MVN86, MC89, MMK98, PR96, RT00, RM97b, TW07]. Stade [FM99]. Staffelstein [GS88]. Stakeholder [PM08]. Standard [Bar08b, EKS00, Sch02, TDB^{+}06, Uni81, Bar97, DRS05, FR99b, Rus05, TD97, TDBP01, Küs01]. standards [CT99]. Stanford [BPCS06, Bla03a, DS01a, Dil94, HPV00, KS10b, Sab10, TG02, vEDA04]. star [BJM^{+}05]. Stará [ABH12, KU06]. State [BBBB08, Dad01, DFG^{+}12, DMW06, DRSS01, FGK^{+}10, Geu98, GKOT00, KEDW07, ZT04, ALM95, BFS92, BGR03, Bur97, FW98, God96, MPS93, PGV93, Sif90, PR98a, YJKK06, YJKSW10]. state-explosion [God96]. state-of-the-art [ALM95, MPS93]. Stateless [Sto04]. Static [Le 94, Lis89, AV08, CF99, Cou93b, CS96, Cou01, Cou03, CM11, Gia04, HS05a, HP02b, Le 94, Lev98, MS12c, Myc95, NF07, Pal00b, PS09a, Van97a, Yah11, Yi06]. Station [Rau04a]. Statistical [AB12, ABF^{+}07, CAD^{+}02, CFB11, DF02, DFT04, DFF06, DFS08, DFM11, DFT12, FCD^{+}04, GL10b, GHI^{+}12, HWW^{+}10, Heu04, dVLKR^{+}08, LM08b, Neu04, SGG06, WNL05, Win09, YKF^{+}06, RS96, ABF^{+}07, CPR^{+}10, CKP^{+}12, CKMS05, CRYS09, Mic90, RKS89]. statistics [CKPB07]. STCS [NSM07]. STDBM’99 [BJS99]. Steady [Dad01]. steam [ABL96]. STEG [LCY^{+}11]. Stein [Dit88]. Stein-Ebernburg [Dit88]. Stellenbosch [BL10a]. steps [MvOvRdV05]. stepwise [MO97, dBdRR90]. stereo [Sch99]. Steyr [CCP96]. STIA [DFGN12]. Stirin [Kra99]. Stirling [Koz04, RMS04]. STM [CLBP11, MFG12]. Stochastic [ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12, BHH^{+}04a, Dad01, HT06b, SBH07, SBH09, Wol07, AS03, Dob81, HKNW07, LKZCS05, Ste01, WZ09]. Stockholm [Alb99, And95, ABJ02, AI98, AM02, BCS09, BS10, BP12b, Dam97, DGT10c, Dol06, Ekl94a, Ekl94b, EHSH92, Gal97a, GR12, GH12b, HAM95, KS03b, MR91, MSY00, Mug97, Sir12, SSB90, WB00]. Stony [Cle92]. stop [Pfi96]. storage [Ful75]. Storrs [BGPR10, JW05]. Story [WPR08]. Storytelling [ALL^{+}10, BST03, CD07, GSH^{+}04, GMI06, IZP09, STA^{+}11, SS08b, Sub05, BST01]. Strasbourg [CML^{+}10, Des86, NS87, RS01, Sub05]. Strategies [BEM91, BKN07, FR00, FRS03, FS07d, FS08, FS09b, FS10, Sch98c, FR95a, FR96, FR97, FR98, FR01a, FRS02, FRS05, FFRS05, FNPW05, RF99b, CS87]. Strategy [FS07a]. Strbske [GC81]. Stream [RB08, DXS91]. strength [Waś96, WL93]. strengthened [FJL97]. Stresa [ADCR89]. string [Kli85, ATM09, AM04, CBGCCZ12, CL10, CN05, CFS06, GSS11, KTH09, LO02, NdMO03, ZBY07]. Stroudsburg [LPH09, MGH09]. Structural [Geu98, RA00, PpWR96, RTB02, Smi98, CAD^{+}02, EH12, FG06, FCD^{+}04, GHI^{+}12, HWW^{+}10, KY11, KS04c, KŽ10b, dVLKR^{+}08, PSE10, PR05, PZ07b, SF08b, YKF^{+}06]. Structure [GKS12, GI03, Hof03a, Lei74, Pol01, SGG06, KvG98, MB96, OFK05, Sel86, Roy87, Sch86]. Structured [CM00, Bil97, FR95b, Fer96, Fer98]. Structures [AI06, Ama08a, APHA11, Bir12, CFL10, CRW^{+}12, DST01, DSS03, DSZ07, DGST09, DIS11, EH08, GG98, HS74, Hof11, Jür02, Kai11, Kuh10, MMS93, MC98, Ong10, PPH07, Pup10, RDK09, Sas05, SHØ06, Sei07, Wal04b, WPD04, dA09, dMLG03, dSOSR04, dSOSR06, Akl95, CT99, DSS89, DSS91, Deh93, Deh97, Deh99, DLOS05, DS01a, EEM96, Ell95, GM00, Gor03, Gre75, Gun88b, Haj96, HM01, Luk97, MO92a, Mul93, NE02, Niv98, Oss99, PN93, Pol01, PCA02, SDM05, TDS94, Tho99, Tiu00, Yao82, YK82, vM94, MRS97, Ove83, Wan93].

Struk-

tur [Lei74]. Student [IM10, LS12a]. Studies [Ber08, CS02a, CW03, FBHW00, Ham00, KL05, KW07b, LM06a, MS12b, McC96, MN98, Rei99, SM01, TG04, XCd03, BB96, FKS86, Lem97, TPB79, TBG05, Gol11b, Lam96, vdADO00, WSL93]. Study [KMK10, Hil78, Kli85, LL95, RR89, BMS96, Lin03]. Stuttgart [DD05, DN09, KGRD10, RR03, RH98, RFBD09]. style [AH97b]. sublogarithmic [Sze94]. subroutines [ZWS81]. Subsequences [GGHS07]. Subspace [SGG06]. Substantial [NM08]. Success [WPR08]. Sudan [Gol11b]. Suhl [AJM87]. sul [AMT03, AMNT06, AHJ07a, AHJ07b, AHJ07c, EFZKM07, EHMM09, FSV09, GG98, MT02a, MTV09]. SUM [BG11a, DH10a, GP09b, GL08, HLFS12, PS07b]. Summarizing [DLN10]. summation [Ger04]. Summer [ACP01, AAB^{+}07, ABW10, BBF^{+}06, BBM^{+}08, BDPS02, BCSW10, BC08, BBPP09, BHK01, Cas01, CSW06, CMT01, CFR92, DF09, EM05, EK12, ÉEMS04, FLVS11, GG98, GI03, HMT99, Hor06, HPSZ08, HPZ10, LSV06, LVS08, LMŠT01, MSY08, MS03c, MN12a, MN12b, Mül10, OS09, PdA^{+}11, PvS07, RG03, TFE^{+}09, Vää99, ZHP12, AM91c, BCG02, BS76, BB79, BvLR04, MST92, MSS05b, MO92b, Paz97, TBG05]. Summer-School [ACP01]. Summit [BFT^{+}08, DJ08, LMSS07, LCDT08, LDC^{+}09]. Sunderland [JNS05a, MT04b]. SUNSET [CGMP11]. SuperComputing [DHO08, DO09, DO11, HPP88]. Supervised [ST12b]. Supply [LY01]. Support [AM03, CMSM^{+}10, Kam99, KCO04, LV02, MLK04, MGSM12, NM08, SBC^{+}94, TKSN11, Bar03, Bre97, Bui87, Her92b, Ish99a, LG95, PS97a, Ran93]. Supported [SCL^{+}06, SLL^{+}07, SYY^{+}08, SLB05]. Supporting [MS08a]. supportive [Mue01]. Surathkal [TPCB12]. surface [Cip95, dB93]. Surfaces [Ahn04, BCC^{+}12, DFL^{+}10, HMS09, MSW07, WM03, MBS05a]. Surgery [TGM97, AD03]. survey [Bar85, BKL^{+}91, MT99c, Pup99, Hir81]. surveys [Fre98, Gol11a]. Susono [HW06a]. Sustainable [CS12c, LFJI10a, LFJI10b, LJS^{+}10, FS07b]. Sustaining [ZRP^{+}11a, ZRP^{+}11b, HR08, WKS^{+}10]. Suzhou [CLZ^{+}09, LFP^{+}09, PMC06]. SVI [MDH^{+}11b, MDH^{+}11c]. SVM [LV02]. Swansea [BBLT06, BBH01, BGM94, FHRR05, Kul09]. Swarm [DGB^{+}06, DBB^{+}08b, PDSD10, TST10a, TST10b, TSCW11a, TSCW11b, TSJ12a, TSJ12b, DBB^{+}04, DBD^{+}10b, DBB^{+}12b, PDSD10, PSDS11a, PSDS11b, PDSN12, RKS^{+}12c, ŞS05a, ŞSW07]. SWAT [AF06, FK12b, Gud08, Kap10, AI98, GK90, HK04, Hal00, KL88, KL96, NU92, PS02b, SS94]. SWDB [BTF05, CCG08]. SWDB-ODBIS [CCG08]. Sweden [Alb99, And95, ABJ02, AI98, AM02, Asp98, AAR05, AS07c, BR12a, BCS09, BS01a, BG03b, Bv97, BNS00, BH02, BP12b, Bru04, CDNS00, Dam97, DGT10c, Dol06, DNS95, EDGP11, Ekl94a, Ekl94b, EHSH92, Fri93, Gal97a, GB97, GR12, GK90, GHC^{+}06, GH12b, GGK^{+}09a, GGK^{+}09b, Gud08, Gyi96, HNH^{+}09, HAM95, HK11, HCF06, HCR06, JM95, Jeu98, Joh03, JP94, JP96, JC12, Kåg98, KEDW07, KL88, KS03b, Kar83, Kir96, KSW99, LK10, LKC93, LN03b, LR94, MSB12, MR91, MSY00, Mug97, Nie96, NR08, OKBMB10, PY05b, Pet92, PVC^{+}99, Pol99a, Sir12, SSB90, SDK^{+}08, TSKGB04, WB00]. Switching [MSS05b]. Switzerland [AK07, AP09, AHS11a, AHS11b, Ame91a, AJ11, AE06, BCT02a, BB05a, BF09, BN91a, BCB97, BV11, BL06b, Bor92, BCF06, BM11b, BRR97, BM06, Bur90, CV11, CC04, CK12a, CM95, CG09, CFR92, DGN01, DGJ09, Dub09, DH07b, ES10, EK07, FRvKS10, FPFR05, FR97, FLF^{+}08, FMN99, FK95b, GGHN97, GW11a, GKW02, GB07, Gün88a, GS91, Gun94, GKOT00, GM07b, Gut94, GW00, HS07, HS06, HKNW07, HKV10, IMW04, JS97, Jen03, JNS05b, KTC^{+}05, KLPP03, KM00, KP10b, KWT03, KF12, KBD87, LV05, LTT07, MTT98, MT09b, MMBND07, MN02b, MJ07, MW08, Mic10, MT05b, NJ10, Nie94, NG10, Nyb08, Pai09, Pla96, RG05, RMB10, RGW01, RLT^{+}08, RKM06, Sch93b, SF08b, SMPU98, SF96, Smi06, SH09, SS99, STTP02, SSKR02]. Switzerland [SD10, Str96, TMG10, TWN92, Tro11, VGP08, Vau05, Vck99, VDÉ^{+}12b, VDÉ^{+}12a, WT96, WMR00, WVD02, Wes01, WNE99, WCJS10, WZA10, WD00, Zit01, vG02]. SWSWPC [CS05a]. SWWA [MT^{+}03]. SWWS [MTH05, MTH06a, MTH06b, MTH07a, MTH07b, MTH08, MHD09, MDH10c, MDH11a]. Sydney [Ami98, AJS00, ABD08, BTBP^{+}10a, BTBP^{+}10b, BKM08a, BHL11, CMS08a, DSZ04, DER94, Du00, EEM96, Ekl04, FL09, FK02, Foo99, GMV12, HNQ08, IPRO08, Joh97, KS02d, Kob06, LG05, Man09, MS03b, MB09, PKSS02, SNY06, SP90, SA96, SH10, Stu08, Tob80, VPM97, VM99, VM01, WPV04, WMB10b, WMB10a, ZJ05]. Symbol [VSTN06]. Symbolic [CBC^{+}02, CIW06, Fit84, Ng79, BC04b, CC93, CC95, CCP96, CL96, CP98a, CRL01, FS03, Fit93, Ger04, HI93, IHQ96, Mio90, Mio93, RS96, BB01a, CKM93, FK95b, Gia89, God05b, HP99c, KS91, Liu11, Mel07, Ng79, NZ03, SC09b, WL03]. Symbols [MÁ05b]. Symeon [KR11]. Symposia [Cha09, DS11b, Gho10, Gie08, HKS09, Kie12, Küh07, RKS^{+}12c, Got83, GAY86]. Symposium [AI97, AM10, ARSB09, AN97, AR99, Agu12, AdBB12, AS03, AV08, Alp11, AF02, AH03, AYBH^{+}06, ATM09, AP10, AMRŚ08, AH97a, AT02b, AGM03, AS07a, AÖ12, AS07b, AHW07, AN07, AN09, Asa06, AiNOW11, AH07, AH10, AD11, ACR^{+}08, ACD^{+}10, ADK04, AE06, BN09a, BYGP95, BSC12, BBB^{+}07a, BBH^{+}10, B^{+}12a, BF08, Bar11, BLS12, BGP08, BGP11, BL04a, BCM08, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BBP^{+}12a, BBP^{+}12b, Bec75, Bec79, Bel03, BMR^{+}04, BHRU09, BE08, BC10a, BGPTF10, BGF06, BSTL07, BU94, Ber96, BRL08, BLSS11, BK05b, BG12b, Bil97, Bis88, BL07a, BR02, BHL90, BH77, Bla90, BMSW12, BKV10, BHHJ11, BBR07, BHR84, BK96, BG07a, BG08, BGPR10, BS93, BGL09, BM02b]. Symposium [Bos00a, BC10c, BS01b, BL07b, BVNW87, Bra96, BC11c, BAH09, BH00, BCH^{+}05, BD04, BS12, BW92, BP93, Bru04, BSL^{+}09, Buc90, Bue04, Buh98, BH11, BDLMM12, BDFS04, BJ01d, BS11f, But10, BKO03, BKO04, BFKO06, BFK^{+}07, BFK^{+}08, BFC^{+}09, CLKL09, CHKL10, CBGCCZ12, CL96, CDCW09, COT98, CP10b, CT06, CFM09, Cas09a, CCK^{+}08, CM12, CL06a, CC91, CSR08, CL10, CMZL04, CPZ07, CWZ11b, CFLR12, CCP10a, CCP10b, Chi04, Chi03, CBS^{+}11, CC97a, CL90, CQSS09, CW00, CCAS12, CI98, CK84, CP99, CMM93, CGM95, CAB10, CÓ11, CGS02, CW88, CHK06, CF99, dRCVT11, CDFP11, CM04b, CS96, Cou01, Cou03, CM11, CMM01, CSSW04, CVÉ07, CMS08b, Cup00, CE00, DW03a, DO02, DF01a, DG06b, DFUR11, DSS11, De 07b, De 10b, DB04, DHRT10, DPV11, Deg03, DGST09]. Symposium [DIS11, DH11, DT89, DNS09, DEQ92, DCM82, Di 97, DPT12, DS96, DVM^{+}11, DH04, DVV07, DR02b, Dit01, Dji89, Dol06, DCFY10, DFT09, DDI09, DGJ09, Dro08, DZ94, DS06, DT06b, ET01a, ES02, EFW93, EMW94, EG10, EF12, EL02a, ENO12, EWZ11, EN74, ERMS06, Est99, FC98, FB12, FL08, Fer98, FR01b, FMS10, Fes10, FKK04, FK02, FJ92, FHH11, FHW12, Fit84, Fit93, FJL97, FHT05, FK12b, FM84, FYM12, Fos99, FHP03, FILP06, FMNS06, FdS12, FN12a, Fre01, FMRW09, FM96, FMY04, GBH11, Gan88, GG09, GH08, GW96, Gel99, GN12b, GS09a, GM11a, GM12, Gia89, Gie10, GS09b, GH12a, GHK97, GL74a, GL74b, GRK^{+}10, GA09, Gol94, GKW02, GMS06, Gom01, GP12, GK02, Gor10, GHC^{+}06, GHP09]. Symposium [GZ06, GHH06b, GPT10, GSS11, GRP10, Gru77, GP09c, GPP09, GS91, GY03, GYD^{+}06, GPL99, HNH^{+}09, HRZK02, Hac75, Had03, HM11b, HM94, HW06a, HSJ07, HKB99, Han98, HMT10b, Han07, Han09, HS06, HP95, HK12, HKI08, HS76, HK92, Hen06, HP94, HS95a, HP02b, HPP06, Hic04, HSS07, Hil06, HRSS08, HKLP12, HM96, HK10b, Ho12a, Ho12b, HKT12, Hol07, Hol08, HM09, HNF08, HNQ08, HTMC09, HZB07, HL91, HM11c, HRA02, Hu09, HH03, HW08b, HMCL06, Iba92, IKO03, ICSY07, JOD98, JL08, Jan88, Jay99, Jay04, JS97, JNHJ10a, JNHJ10b, Jen01, JCC^{+}12, JSK10, Jha10, Jon90, JP96, JVA04, Jos00, dCJB12, Kah79, KMP84, KN09a, KS04a, KLZ07a, KCYL08, Kap10, Kap08, KS12a, KTH09, KKMW07, KW07a, Kei09, KT06, Kin08, KJB09, Kla12, KP01a]. Symposium [Kob06, KKM^{+}11, KH02, KW07c, KKS10, KM00, KYT01, KU06, KN09b, Kra99, KB92, KR02c, KMVY06, KMYJ08, KLC07, KRSR11, KK07, KS96, KU01, KU09, KR04b, KV11, KS08b, KH09, Kut98, KPW99, KCG09, LJS08, LBNF08, LO02, LdRV94, Le 94, Le 02, LMW01, LYMM05, LSL^{+}09, LYBU10, LHH10, LIJ97, Lev98, LSY^{+}06, LV06, LPH09, LLM^{+}08, LN03b, LP10b, Lio04, LKT08, LR05a, LCB97, LFH^{+}07a, LFH^{+}07b, LFH^{+}07c, LR09, LZP^{+}11a, LZP^{+}11b, LZP^{+}11c, LW12, LO10, LS06b, LM98c, LS12b, LV07b, LS10b, MZ07a, MCA08, MB82, MCKB06, Mag98, MMMP12, MRK05, MNS05, MRvM07, MW91, MSP^{+}09, MZ07b, MSZ08, MNZ09, Man75, MSS12, MS06a, MS10b, MS10c, MS12b, MSZ90, MEAC04, Mar01, MR12a, MRZ09, MWZ10, MT07a, MTO99, MMR91, MP09, MP95a, MSAdAS05]. Symposium [Maz76, MEB05, Meh85, MT99a, MRSV07, MR07a, MF11c, MCSN10, MMR93, MK03a, MS99b, MR07b, MS12c, Mio90, Mio93, MNS06, MS00c, MVN86, MC89, MS08b, MSB07, MM97, MS11a, MMK98, MNTY05, MS11b, MB97, MJL02, MZWZ05, Myc95, NDB94, NYHO07, NMPM08, NdMO03, NT92, NM94, Nes95, Ng79, Ng93, NLKJ06, NGHJ07, NJHJ08, Nic86, NR00, Nie96, NF07, NL99, NY93, Nor97, Nür03, OT08, Oho03, Oka03, OZ01, OMMSP06, OW09, OPS11, OA93, OST11, Pac05, PHZ11, PGM98, Pal00b, PS09a, PZK07, PR11, PB07, Pat02, Pat00, PR84, PT08a, Pel07, Pel11, PP12b, PS96a, PRT06, PBH03, Per11c, PTM^{+}11, PPBL12, PRdRW09, Pol97, Pol99b, PT91a, PT91b, PRR94, PR97, Pue07, PR96, Qin10, Qui98, RF01, RCC12, Ram01, Ram08, RS10, RUL04]. Symposium [RRU05, RS00, RZ91, RK93, RZ94, RS99b, RO00a, RRBE09, RR98a, RT00, RM97b, RS12, Rig08, Rob74, Rob80, RW86, RL11b, Rol98, Rol99, RV03, RSW12a, RZ12, Sag05, SW07, SS09a, SS11c, San01, San94, Sav06, ST00, Sch04b, SCHS07, SC11, ST12a, SNZ08, SVV12, Sei12, STT04, Ser12, Ses06, SVA05, SPK01, Sha07, SXC04, SCR06, SM11b, Sim92, SMEGA10, SS07c, SS07b, SS09c, SS11d, Smo00, SG03, SPO11, SBM11, SMF^{+}11, Spi95, Ste01, STY^{+}07, Stu90, SZT^{+}08a, SZT^{+}08b, Swi99, TFK11, THL11, Tar91, Tau08, TBKO10, TYZ11, TDG91, TW12, TG00, TW07, TSS11, Tok07, TP09, Tou94, Tou96, TNGK08, Ued10, Vah09, VJKT00, Van97a, VJAA03, Vid12, VKJ03, Vyt91, WHL09, WC95, WLY05a, WLY05b, WLY05c, WYZ06a, WYZ^{+}06b, WYP12a, WYP12b, WT07, WZ09, WVD02]. Symposium [Wes01, Whi98, WH95, Wii04, Wii05a, Wii05b, WR09, Win78, WHR10, WL93, XLKZ12, XZCL07, Yah11, Yan11b, YŞ03, Yi06, YWG04a, YWG04b, YGGÖ05, YKM06, YKMJ09, YAKMJ10, YAKMJ11, YAKM^{+}12, YHZ09a, YHZ09b, YHZ09c, ZK06, ZdC08, ZLK10a, ZLK10b, ZRTS03, ZJS^{+}02, ZBY07, ZBG98, ZSP10, ZQ98, dBM10a, dBM10b, dBBGdR03, dBBG04, dBBGdR06, dBBGdR07, dBBGdR08, dBBM09, dBBHL10, dO98, dRLP98, dSK12, dSTP11, tCZ07, vEDA04, vG02, vKS12, vdPS08, ACvLdR03, AJS05, AH94, AR04, ABD04, AL01, Apo92, Apo93, ACP05, Asa96, BYCC03, BMSMS04, BBKP06, BGD^{+}05, Bil98, Bjo87, BZ83, Bli75, BMJ88, Boh75, BL05c, BW97, BFKO05, CYGL04, Che03a, CJK88, Coh96, CG94, DDFM05, DS05b, Dem80, DD06, DZ03, DD05, DH06, DGR92, EH95, EW91, Eta05, FKP^{+}05, FC04, FS09a, FT04a]. symposium [GU95, Gia04, GS00a, GPV00, GC81, GRW86, HMRT05, HM08, HS05a, HN06, HCS^{+}05, HT05, HC05, HKNW07, JS05, Jos88, KG74, KZ97, KSH05, Len93, LMV98, LKZCS05, MT89, Mey01, MT86, MR02, MO92b, Nes99, Nic75, NS81, PCG^{+}05a, PC99, PT05, Raj02, RM96, RS97b, RBGS04, SMD04, SRG04, SV97, Sko85, Sta95, Tob80, VZ06, Yao82, Yi05, ZHF04, ZS05b, dBdRR94, dBBGdR05, dSG05, dLWC03, vL94a]. symptoms [Nok91]. synchronous [Lis89]. Synchrony [ECV^{+}10]. Synergies [TK05]. Synergy [CDS09]. Syntactic [CAD^{+}02, FCD^{+}04, GHI^{+}12, HWW^{+}10, dVLKR^{+}08, YKF^{+}06]. syntactical [PpWR96]. Syntax [MR12b, AF99a, MD96]. Synthesis [Alp11, Bru04, De 10b, Han09, Hil06, Kin08, Pet02, Pue07, Sch03b, Vid12, ZJL06, BJ86a, Bos00b, Eta05, Fle99, FT94, Fuc98, Gal97a, HS94, Lau01, LB90, Leu03, Pro96, Sch99, Sha04]. Synthesizing [Lis89]. synthetic [TP97]. Syracuse [MMO^{+}11a, MMO^{+}11b]. Syros [DVVA08]. System [ADM^{+}12, AGCP^{+}11, AS07c, BEH^{+}85, BLKP08, BB89, BHL11, BvdHPP09, BZ03, CKK^{+}09, CW07, CM03b, CvRVSN12, Dau03, DW10, FTAZ06, Fri06, KO11, LFIM07, LLIH07, LFJI10a, LFJI10b, LJS^{+}10, MPK04, MvL10, OO12, PMKG07, SKMT04, Sch01b, SBC^{+}94, SO99, SKM07, SM09, VAM06, XBM12, vLS11, BGW93, BKL^{+}91, BBP96, BBZ00, BBZ01, Bui87, EKL^{+}05, FHM82, GB99, Gut94, Her92b, LDTU04, LG95, NY83, PVV05, Ram98, Ran93, RR03, SS91, TG12, AW05, Est99, GR06b, HM87, KH11c, Mag98, Nag01, HKB93, Kis95, Kum90, Stu85]. systematic [Lem97, US90]. Systems [AAW04, ARS09, AL11a, AKK04, APR^{+}06, Abr07, ADM11, ABDW10, AKM10a, AKM10b, AE05, AD06, ACJ09, AR10, AP04a, ACK00, AMRŚ08, AFGvE10, AFGvE11, APW06, ABJ02, AH98, ADMH04, ADM^{+}08, AMCL02, AYB^{+}09, ATO^{+}09, ALR11, AMNT06, AHJ07a, AHJ07b, AHJ07c, AB06, APP12, ASAM07, ACJ08, AJN12, Avg12, BBMB^{+}08, BBD^{+}10, BA06, BHH^{+}04a, BMS^{+}08a, BBT08, BMU^{+}09, BdB08, Bar11, BLS12, BG09, BPR05, BCL03, BL05b, BPR02, BL08, BBB07b, BCG^{+}07, BDG04, BKR09, BLJ08, BDW08, BMSHW09, BFBS11, BI03, BH07, BDZ08, BPZ09, BI11, BCP12, BC06, BDSW09, BRL08, BG11b, Bha00, Bha05, Bha07, BKW08, BKW09a, BB03, Bis88, BTPPS06, BTPPS07, BTBP^{+}08, BTPPS09, BTBP^{+}10a, BTBP^{+}10b, BTKP^{+}11, BTPP^{+}12, BMO02, BPS06, BRU06a, BHU03, BPD06, BJ07]. Systems [BDDS05, BDDS06, BDDS07, Bou11, BMYZ91, BMYZ95, BIMW12, BAP^{+}05, BKL^{+}12c, BBT10, Bra80, BKZ05, BJ01c, BGHS05, BKW06, BHQ10a, BHQ10b, BUHS08, BGR08, BJK^{+}05, BSH06, BHJY07, BB02, BD11, BK04b, BSS00, BBB06, BHH08, BLVV07, BF04, BkHJ^{+}04, BSG03, BSPB07, CPS05, CSPC09, CG07, CJ11, CG12, CL08a, CT02, CG04, CT06, CFM09, CPSA10, CdLG10, CJ08a, Cas09a, CIT11, CA08, CD00a, CGP02, CCPP12, aCF06, CH10, Che99a, CH04, CLLD09, CY10, CTS10, CY11a, CFLR12, CYZ12, CLZ06, CEQ^{+}11, CHCA09, Chi04, CBS^{+}11, CPC07, CGG^{+}07, CN06, CHAL04, CDY^{+}11, CHA03b, CB06, CVV12, CDE^{+}07, CG09, CGS07, CGK^{+}12, CF09b, Col07, CDN12, CAP08, CWO^{+}09, CRS10, CKW11a, CKW11b, CSA^{+}12a, CSA^{+}12b, CCSW12, CKTW12, CDT11, COB^{+}08, DLFP02, DYCD09]. Systems [DO02, DDE04, DSLT08, DSRW08, DSLT10, DSHL11, DG06b, De 07b, DKK08, DFPV11, DG02, DGJ^{+}05, DPV11, DG09a, DT89, DBB12a, DV07a, DLW12, DPT12, DPA07, DD10, DFN10, DLGJ10, DPT08, DB07, DCFY10, DP97, DMV01, Dro08, DKR02, DP06, Dwa00, EHKP05, EBT11, EM08, ERRW03, ES02, EKR10, EM06, EK10, EOSW07, EBT07, ENO12, EWZ11, EIRW06, EM10, ERMS06, EKM06, ED06b, FT11, FPRAW04, FR11, FKK09, FWZH06, FKB02, FFM05, Fis08, FST09, FvdTDG12, FK12a, FLUV02, FP08, FKBA12, FSP09, FLT07, FMY04, GHJ06a, GHJ06b, GHJ06c, Gae96, GTT04, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, GVCA07, GLZ^{+}03, GCL^{+}06, GVT08, GNR07, GS09a, GGS09, GAA06, Gie10, GKL^{+}11, GR12, GH06, GHSW04, GM06a, GLC^{+}11a, GLC^{+}11b, Gor10, GZSC07, GW06a, GH12b, GM06b, GP08b]. Systems [GSW06, GAM^{+}11, GH04, GH07b, GMV09b, GKMN10, Gue01, GP09c, GHP^{+}08, HNH^{+}09, HRZK02, HKR07, HG02, HPTV05, Ham08, HT09, HKW09, HMT10a, HKW^{+}10, HLSZ11a, HLSZ11b, HT11, HKW11a, HKW^{+}11b, HHMT12, HKW12a, HKW^{+}12b, HTW02, Han02, HKP08, HMTH10, HKI08, HZK08, HZ10, HU08, HA02, HRB12, HP06, HC11, HT06a, Hig05, HRT^{+}03, HRR^{+}06, HPS09, HR08, HL10, HKP11, HRM^{+}10, HMMW10, HSH08, Hu09, HMBD09, Hug03, HKH10, HS08c, HHG11, HK07, IKR00, ICSY07, IB07, IAC06, IR07, IST07, INR07, IGN05, JM05, JM11a, JNHJ10a, JNHJ10b, JP04a, JCC^{+}12, JKN^{+}12, JM11b, JDAW12, JLW07a, JJF03, JMK09, JC12, Jos00, JQ04, JN12, KS03a, KMTW03, KKP02, KKGK11, KLZ07b, KR10a, KSS08, KP04, KLM^{+}07, KMF06, KSB10, KMSB11]. Systems [Kim05b, KM04, KMPH09, KILW10a, KILW10b, KLLP12, Kob06, KMS06, KBHSW06, KHSM^{+}08, KW07c, KHW03, KDH^{+}11a, KDH^{+}11b, KDH^{+}11c, KDH^{+}11d, KLPP03, KFMM06, KS07a, KS07b, KKMN07, KLT^{+}07, KP09, KR10b, KS00, Kra03, KSHM09, KOS07, KS12d, KPRS07, KRSR11, KH12, KS08b, KMR12, KL02, LJS08, LSW03, LN04a, LPOA^{+}06, Lau12, Lau00, LB03, LLWL04, LTW06, LKK^{+}07, LLPH09, LN09, LMMN10, gLPZ^{+}12a, gLPZ^{+}12b, Lei74, LT05, LTÅ^{+}11, LVP04, LM08a, LST^{+}05, LHDD08, LSTZ12a, LSTZ12b, Lin01, LMP04, LP10b, LNS11, LBDW09, LLWL09, LW12, LPA00, LHK02, LHJ08a, LHJ08b, LHJ08c, LSU07, LMM10, LGR^{+}04, LS12b, LS09, LTT07, LCDT08, MT09c, MS10a, MN02a, MPS06, MSS11, MPdSM07, MRŠ03, MMV04, MVC07, MSZ09, MVL11, MBH^{+}09, MRZ09, MWZ10, MT07a, MPR07, MRPM10, MRP11, MPR12, MG07a, MTP^{+}02]. Systems [MTS^{+}03, MT^{+}03, MT04c, MT04d, MTC04, MT06a, MT06b, MTH06a, MTH06b, MT07b, MT07c, MTH07a, MTH07b, MT08a, MT08b, MTH08, MDH09a, MDH09b, MHD09, MDH10a, MDH10b, MDH10c, MDH11a, MDH^{+}11b, MDH^{+}11c, MMO^{+}11a, MMO^{+}11b, Mei08, MT08c, MCA^{+}07, MCSN10, MB09, MM04, MPPU11, MGH09, MMM05, MDP04, MDPQA09, MDPQA12a, MDPQA12b, MMC05, MBJ^{+}07, MHW12, MR11, MCER10, MSUB04, MSKY10, MMM11, MNTY05, NNS03, NPPDG06, NP11a, NHJ04a, NHJ04b, NHJ04c, NKPH08, NKN^{+}05, NGHJ07, NJHJ08, NBGA08, NKC09, NLŚ10a, NLŚ10b, NKJ11a, NKJ11b, NCBT04, NWHR06, NCJK08, NVSB^{+}07, NBM09, O’H98, ONC^{+}11, ONPR07, Oho03, OA07, OS03a, OYA04, OR08, OV09, PAV^{+}10, PV07, PS02a, PHJ03b, PHJ03a, PVW10, PR07, PT09, PPCS12, PCN12a, PCN12b, PCN12c, PA01, PKM^{+}10]. Systems [PH04, PMMR06, PRK12, PV02, Per10b, PS04b, PDF^{+}09, PVTG07, PSM10, PRH10a, PRH10b, PY05b, PBSB09, PMWO02, PVV10, PS09b, PG09, PISW05, PT11, PLMV12, Pri03b, PMGO05, Pri05, PZ05, Pri06, PCE06, PHPL06, PP06, PIMN06, Pri07, Pri08, PDAN08, PBP09, PBG^{+}10, PBPdV11, PPdV12, PRR05, PS12c, QLXL07, QSBC98, QBY^{+}12, RPR08, RM09, RWSS12, RMDS10, RFBW12, RR08, Ram08, RUL04, RRU05, RA12, RS05, RO00a, RT07, RRBE09, RA00, RSS06, RS12, RK12, RR04b, RK02, RKW04, RCS10, RG11, SPG10, SS08a, SNY06, SW04, SS11b, STP^{+}10, SKR11, SMBQ07, Sch10a, Sch04a, SUW02, Sch04b, Sch12, SF11, SvS00, SS05b, SSN08, SS09b, Sco02, SVS08, Sei12, STT04, SP08, SO09, SKRZ04, Ses06, SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, Sha07, SS06b, Shu11b]. Systems [Shv06, SPON08, SH09, SXC05, SS99, SvdMOP06, SDH03, SS06e, SP12, SHYEF08, SHUH08, SC00, TBE08, TFK11, TTM10, TFE^{+}09, TDG91, TKSN11, TGH06, TBH03, TSS11, TQYX11, TT06b, TTF07, Tro11, THR03, TtHVS06, TNGK08, Ued10, UBM02, UDF06, VY93, VAR12, VWH06, VBH07, VRHJ09a, VRHJ09b, VG12, VC03, VV07, VS05, VLY09, VASP09, WAS06, WRP07, WKP03, WJ05, Wan05a, WJS^{+}06, WGLL10, WZJX12, WT07, WPE05, WHG07, WPM06, WPM07, WBD08, WCK11, WWD99a, WWD99b, WKB04, WBHR08, WH11, WZ11, WANL08, WCGB05, WKD08, XYZU11, Yak00, Yak02, Yak04, YN06, YYI^{+}09, Yan11b, YX04, YMY^{+}10, YKM06, YKU11a, YKU11b, YYH^{+}12, ZPT^{+}06, ZZ04b, ZZP^{+}10, ZHLW12, ZRTS03, ZSP^{+}04, ZYDL09, dCVK07, dFEN04, dLGR05, dLGR07, dLDG^{+}08, dLFG^{+}09, dMS06, vEGW09, vM94, AM91c, ASP85, AKK03]. systems [AHS96, Ame91b, ABS91, ADW05, And95, ACL98, Ant95, Ant97, AKL^{+}99, Ari02, Arm02, ABGP05, Aus93, AAN00, BJM^{+}05, BDGM05, BK05a, Bat01, BSG84, BB95, BR92, BCST99, BC04a, BB05b, BU94, BR97, Bha95, Bha02, BJ86a, BdBK03, BMS95, BTPPS05, BH77, BZ83, Bla90, BU01, BP95, BS05, BKM01, BL06c, BW05b, Bri95, Bri97, BMS96, BR98, BSA05, BKS96, BS06c, BBB98, BMP96, CL96, CC05, CE01, CW94a, CR05a, CRW97, CCL02, Cha03a, CKO01, CGLR05, CH99, CNY95, Cle99, CMV96, Cro03, DS05a, DH01, DM01a, DN04, Deg03, Den85, DTV02, DJL88, DL95a, DSV01, DOS99, DPA05, Dit88, DGN01, DL04b, DH06, DGR92, DN93, DL81, DKN02, ERW99, EM03, Enc80, EW03, ES00b, Eve75, FV01, FV03, FV04]. systems [FV05, FRA05a, Fit93, FP05, FGM92, FGL96, FW92, GH03, GFV00, GHR89, GB97, God96, GP98, Goe98, GLS05, GN90, GS00b, GSB78, Gra05, GL92, Gue90, Gun94, HMRT05, Had96, HZ05, Hal96, HI93, HT97, HCT02, Han98, HWP01, HS76, HNR05, HS98a, HT05, HML97, HRTR05, HW03, HH03, HS99, ILR95, Ima99, Ish99a, IH90, IHQ96, JPBT05, JOD98, JL95, JO99, JE06, Joh01, JP96, Jos88, KG74, KJ88, KO91, Kar91, Kar94, KN91, Kat99, KS02c, KD05, KHJ05d, KDS01, KR00a, KR00b, KR01, KN83, KP99, KP01b, KNP01, KRMM^{+}02, KS05b, KM00, Koy92, KMO^{+}05, KS97, KKA05, KBD87, KA05, KBC05, LY04, LM98a, LSDP81, LPSD83, LdRV94, LSW00, LSW01, LSW02, LT01, Le 02, wLN99, LL95, LV05, LPH00, Lis89, LMV98]. systems [Liu01, LB78, Lou92, LMŠT01, LK00, MB82, MML97, Mal97a, MP03a, MS96a, MY01, MLW93, MŠKL02, MMP03, MPB05, MD04, MR91, May01, MT05a, MTH05, Mil00, Mio90, Mio93, MdPA98, MVA01, MT05b, MDBF01, MSY00, MB98, MB87, MSS05b, Neh88, NOOS05, Nie96, Nie94, O’D77, Oka03, OP97, OK96, OPKR05, PSP^{+}02, PWZ98, PP01, PZ92, PF05, PT98, PBGK02, PGBK04, PVP05, PMD90, PMD92, PMD94, PMDA96, PMD97, PMDK00, PV04, PT99, Pri03a, Pup99, RMR05, RM97a, RCA90, RZ91, RK93, RZ94, RM96, RS97b, RS99b, Ras01, RR98a, RT95, Rob93, RBC93, RS74, RV98, Rus05, Sag05, San01, San94, Sch96, Sch98a, ST00, SS01, Sch97b, SPZ90, Sel95, She01, SCG98, Sif90, Sin94, SMPU98, SLL01, Smo00, SLPM96, SG96, Ste98, SZ05, SSB90, SW97]. systems [SW05, ŠPTM00, Swi99, TC98, TG02, THT03, VvS99, Vck99, VT97, Vyt91, WT96, WR01, W^{+}12, WB00, WS96, WPM05, WBW94, WR03, X^{+}12b, YBM87, YZZ^{+}05, Yi05, ZL97, ZL98, ZOM05, ZWS81, dBdRR90, dBBC02, dLGR03, dLGR04, dPMA98, vSN01, AMZ09, Bai05, BL05b, BKR87, BJRT06, BSG03, Cai05, CdLG^{+}09, EIRW06, GTP12, Geu98, HKW11a, IB07, JvdA09, KHJ05a, KHJ05b, KHJ05c, KS00, MdPA98, Nau02, PA01, PPPM10, Pas02, Paz99, PBCF08, PBSB09, Pup10, SKS^{+}09, AB93, AW05, BB96, BK81, Bre97, Cas81, CF80, DC01, Gro93, Hoo91, Ish94, JFM96, Kom97, Kra97, KU94, LSW81, Leu97, Mil80, Nay95, Neb90, OK90, Ooi90, RR93, Rus05, SYBM88, Sch91, Sch97b, WP90, WS00, van97b, von79]. Szeged [CÓ11, CDG89, Esi93, EF03, Ési06, FG95, Gec81, HY03, KNP06]. Szklarska [KPW99].

Tableaux [BM11b, EF02, GW09, MP03b, Oli07, BHP95, Bec05, Dyc00, Gal97b, MMMO96, Mur99, de 98, BHP95, Bec05, Dyc00, EF02, Gal97b, MP03b, MMMO96, Mur99]. TABLEAUX’98 [de 98]. TACAS [AL11a, EM10, FK12a, GH07b, HZ05, HP06, KP09, RR08, Bri95, Bri97, GH03, GS00b, JP04a, KS02c, MS96a, MY01, Ste98]. TACAS’99 [Cle99]. TACS [AI97, HM94, KP01a]. TACS’91 [IM91]. TADA [FS07a, LSJ06]. TADA/AMEC [FS07a]. Taejon [CI98, YKF97]. TAFA [MOT12]. TAGT’98 [EEKR00]. Tahoe [BBKP06, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, HK01]. TAIC [BF10]. Taichung [CM06]. tailoring [KS95]. Tainan [CH04, CHCA09, HXC^{+}08]. TAINN [Sav06]. Taipei [BBDS12, BH11, CBS^{+}12, CHCM11, CLY02, CL06b, CC12, CKLN10, Chi04, CvRVSN12, DHtM^{+}12, DsSL^{+}07, DBB12a, HL91, HK98b, HC09, IY06, KSG^{+}07, Lai03, LT00, LTL^{+}11a, LTL^{+}11b, MPR12, PT05, SSVW12, VSMS12, VKG12, Wan04, Wan05a, WLRL08, YCC^{+}08, Yan11a, YI12, YY01, vdHHsHD06]. Taiwan [BBDS12, BCC^{+}10, BH11, CBS^{+}12, CCL02, CL06a, CHCM11, CLY02, CCH02, CH04, CL06b, CC12, CKLN10, CHCA09, Chi04, CM06, CvRVSN12, DHtM^{+}12, DsSL^{+}07, DBB12a, HK98b, HM11c, HC09, HXC^{+}08, IY06, IAC06, JS11b, KSG^{+}07, Lai03, LT00, LTL^{+}11a, LTL^{+}11b, MPR12, PCN10a, PCN10b, PCN10c, PCN12a, PCN12b, PCN12c, PT05, SSVW12, VSMS12, VKG12, Wan04, Wan05a, WLRL08, YCC^{+}08, Yan11b, Yan11a, YI12, YY01, vdHHsHD06]. Taiyuan [DMLW11a, DMLW11b, DMLW11c, GLC^{+}11a, GLC^{+}11b].

Takamatsu [OBS12]. Takasu [TJHS94]. Talcott [ADM11]. Talence [Coh96]. Tali [Gol11b]. Talking [FNGW08]. talks [PDH99, PDH01, PDHS03]. Tallinn [BGK12, Bir12, BP93, DG12, DV07a, Dur12, EHKP05, FK12a, GL05, KL12, Lau12, MLM90, Nie05, O’B12, Pat11, PS12a, PVTG07, Sei12, dLZ12, HK98a]. TAMC [ADDL08, ACL12, CCL06, yCCZ07, CC09a, KLFK10, OT11]. TAMODIA [CLS07a, EPVW10, FP08, WJP07]. Tampa [JS02]. Tampere [HHLS08, IS12a, IS12b, LS88, OYY03, SMK02]. Tangible [KvETvdH10b, KvETvdH10a]. Tangshan [ZZLL10]. Taormina [BGR03, CFGN11, CGK^{+}12, CCD^{+}12a, CCD^{+}12b, Cra12, HSS96, Pal01]. TAP [BH08, BJ12b, Dub09, FG10, GW11a, GM07b]. TAPAS [MSS11]. TAPSOFT [AM91b, AM91a, BD97b, DO89a, DO89b, Ehr85a, Ehr85b, EEK^{+}87, EKLM87, GJ93, MNS95]. targeting [BR98]. TAROS [GAM^{+}11, HSP^{+}12]. Tarquinia [Par98]. Tarragona [BAH09, DIMV09, DIMV11, DFPS06, HVT12, MVMP^{+}04, MVOF08, TNVDF06]. TARSKI [dSOSR06, MWD06, dS02, dSOSR04]. Tartu [VU05]. Task [FP08, CLS07a, EPVW10, WJP07]. Tasking [SHM^{+}10, BLW87]. Tasks [Hel08, MÁ07a, dPMA98]. Tatranska [Gru77]. Tatras [KN09b]. Tatsunokuchi [DMO98]. TAUES [EYX^{+}05]. Tbilisi [BV02, BLSS11, BGL09, GVM99, tCZ07]. TbiLLC [BLSS11, BGL09, tCZ07]. TC [BBdB12, BPAB07a, BPAB07b, BBK12, BKL^{+}12a, BKL^{+}12b, BPS06, CGJ^{+}11a, CGJ^{+}11b, CGJ^{+}11c, CGJ^{+}11d, CGMP11, CCSW12, CFRR10, DSB10, DLNU11, DC12, DKU05, DPMP^{+}11a, DPMP^{+}11b, EV07, FSTS09, GGK^{+}09a, GGK^{+}09b, HKM^{+}11, JFT11, KD05, KH12, LM06b, MST12, MBVTY11, MNW08, MW08, QBY^{+}12, SH09, SHUH08, SSZ12, Tah09, TEB09, UDF06, LM03]. TC-11 [DKU05, LM06b]. TC-6 [DKU05, LM06b]. TC13 [CP05]. TC2 [MPS93]. TC2/WG [MPS93]. TC3 [SE10, ST12b]. TC6 [ASE^{+}07, BAP^{+}05, COB06, DPLW08, GPAS^{+}06, HDLM09, MCSY07, PVTG07, SMBQ07, TNP^{+}07, BCL03, CGGO03, GCC^{+}02b, MNW01, MKR^{+}04, NdG04, RS00, STTP02, WSS04]. TC6.6 [OS09, PvS07]. TC6/European [Omi00, PPF^{+}00]. TC6/WG6.1 [PVTG07]. TCC [Can08, Cra12, HR06, Ish11, Kil05, Lin14, Mic10, Rei09, Sah13, Vad07, Nao04]. TCGOV [BGPW05]. TCS [AS02a, BBdB12, HMV01, KP00, PR03a, vLWH^{+}00]. TCSP [Tau89]. TD0904 [VEG^{+}11]. TEAA [DW06, DW07b]. Teaching [BCK08, BvDMST11, DB04, GO09, HKV10]. Technical [BI73, BHH74, BSW12, KF74, Lei74, SF74, Sch74, TDB^{+}06, Nok91, Ram98, TDBP01]. Techniques [AB11, AMZ09, ABHT08, ABFH09, ABFK10, ABBFM11, ABFV12, AJRS03, BST03, BS06a, BB95, BJP12, BCI85, Bih03, BCSW10, BK04a, BF10, BKZ05, BD11, BM09b, CC04, CT02, CSW06, CR01a, CJRR07, CP88, CS10, CBS02, CR73a, CR73b, CM09c, Cra05, Dam91, DO02, Dav91, De 95, DV07a, DBSvD09, Dig11, DT07, EDD04, FLVS11, FS07c, Fum97, GP07, GP09a, GP11, GR12, Gil10, GJRR08, GJRR11, GS06, GB07, GQ95, Gün88a, GJRS12, HZ10, Hel94, HKM^{+}11, Jac07d, Jac09d, Jac11b, JM95, Jos00, Jou09, Knu02, KHW03, LSV06, LVS08, LLPH09, LM03, LMT10, LRS98, MTT98, Mau96, MS10d, MLTC11, MNW08, Mid01, MA05c, MK12, NPPDG06, Nao07, Nes95, OYY03, OF15, PA01, Pat11, PV02, PT08b, Pfi01, Pfl02, PJ12, Pre00, QV90, RR04a, Rue93, SZ93]. Techniques [SSJR10, SBM08, Sma08, SS07c, SPS09, SHYEF08, SSZ12, Tis02, Tre09, VW07, Vau06, Vor08, Wan05a, WS06, WPH03, ZZP^{+}10, dFEN04, vHH09, vO04, vdADO00, Bac00, Bar85, BM05b, BCM00, Bla80, Boo91, Bre03, Bre91, BBB98, Cea76a, Cea76b, CJRT05, Com97, Der89, DJRZ06, DJNR09, FS03, FM97, Fia99, FMO05, Fie02, Gac02, Gan96, GJRT01, Gue05, HK94, HBS00, HJRS99, Hsi95, HS98d, JP96, Jos88, Jou85b, KK87, KS03c, Kir93, LY04, LdRV94, LG99, Les87, Mar75, MR97a, NR99, NM97, Nie03, Nip98, Nyb98, Par98, Pic86, PSS98, Ral81, RR98a, Rol97, RV02, ST87, Ste99, Vyt91, Yao82, YK82, CPV03, DKRT06, RDKT01, FP96, Sch97a]. Technological [BBGR03, DGG^{+}11b]. Technologies [AK07, AYBH^{+}06, ABD^{+}12, ADM^{+}06, ADM^{+}08, ABW10, AH10, AKSO06, BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, BBB^{+}07a, BMP00, BTQ03, BBP04, BPW06, Bel03, BMR^{+}04, BHRU09, BLKP08, BCG^{+}11, BPS06, BG07a, BG08, BAP^{+}05, BvdHPP09, BGR08, BP12b, CD07, CWC^{+}07, CHCM11, CLZ^{+}09, Cho02b, CKM^{+}09, CFVV^{+}12, CMS11, DM06, DG06a, DGG^{+}11a, DB09, DS08, Din03, DW10, DGJ09, DFW^{+}11, DHH^{+}08, EK12, EKM06, FSA^{+}07, FKK09, FKB02, FdCLJ^{+}11, FHH11, FHW12, FTAZ06, GA09, GCC^{+}02b, HNH^{+}09, HG02, JNHJ10a, JNHJ10b, JNH11a, JNH11b, JKN^{+}12, KG04, KKW11, KKW12, KO11, KV08, KK09b, LN09, LYBU10, LOST04, LOTY05, LSK08, LS04b, LCDT08, MKPV05, Mal03, Mal07, Mal09, Mal11, MSS12, MS05, MPPU11, MKR^{+}04, NGHJ07, NJHJ08, NHJ12a, NHJ12b, ONC^{+}11, ONPR07]. Technologies [OSV11, PCN10a, PCN10b, PCN10c, PMKG07, PK12, PH06, PdA^{+}11, PBW05, PW07, PW08, RV10, RST09, RV11, RDK09, SNY06, SSVW12, SKMT04, SWJX12, SG07, SBF08, SSKR02, SCKR03, SRT06, SFV^{+}11, Sub05, SV12, TYZ11, TFE^{+}09, VW05, VMB^{+}12, WZH^{+}12, WZH^{+}07, WL04, WM04, XZCL07, YLLZ04, ZPT^{+}06, ZJXC03, vdBK05, Ari02, Asp98, BST01, BMP01, BTQ02, BS02b, BCH^{+}05, CNX05, Chu03, CS01b, DS03, EKL^{+}05, Fed01, FM96, GdlP99, HB97, II00, Kah03, Kam99, KKM04, KYT01, LSS^{+}04, LDTU04, Mal95, Mal97b, Mal99, Mal01, Mal05, MSW05, NY93, PK00, RS03, SLL^{+}06, SILS05, SS99, Str96, TG12, ZTY^{+}05, ZZO03, BS03a, BCF^{+}02, BS06d, CGS01, CJ05, CJ06, GSH^{+}04, GMI06, cHPkC^{+}07, PAD^{+}06, PZE^{+}08, SDH05, KMTU03]. Technology [AT02a, ABK^{+}09, AS02a, AM03, ALPH10, AKG06, AP07, AJN12, AKTT12, BR12a, BEM91, BCP^{+}04, BKLP11, BKLR12, BMN05, BB00a, BC10c, BHLvdP07, BH11, BBC04, Ceb08, CCK^{+}08, CL06a, CKLN10, CDCJL08, CUdlC04, CDS09, DYCD09, DHtM^{+}12, DB98b, DS08, DKW07, ERRW03, Eug09, GTVC06, GBMS11, GZ06, GHI^{+}06, HMV01, HM02, HHJZ12, Ho12a, Ho12b, HNS^{+}12, HJL^{+}09a, HJL^{+}09b, HJBF12, IdKM06, IMW04, IMK06, ISS^{+}06, JSS08, JM07, JM10, JV06, JP11a, JMT11, KP00, KLP10, hKA10, hKhLKŚ10, hKAS^{+}11, KR02b, KGG^{+}11, KS04b, KHL10, KFM07, KT11, LJR^{+}10, Lau12, hLhKcFŚ09, LSL^{+}09, LBM^{+}10, LMMN10, LHŚ11, LHKŚ12, LBSBW98, LLM^{+}08, LMA09, LNW^{+}12, LFK^{+}02, LYhL^{+}01, LWYN09, LR09, LM04, LK07, LK09b, MCK^{+}06, MR07a, MR08, NP09, NP11a, NP12a, NYHO07, NO11, NT06, NLKJ06, NS98, OKHH^{+}08, Oqu03]. Technology [PCK^{+}12, PR03a, PA09, PHOK10, RRR99, RS06, RR12, RLKHL12, RV10, SSO^{+}11, SZC^{+}03, ST02, Sha86, SPSV^{+}04, SKCK07, SC07, SHŚ^{+}07, VW05, VU09, Vet11, VGB11, WHL09, Wan04, WPSY06, WLGB^{+}08, WLP^{+}09, WKS^{+}10, YLW^{+}07, YRWS11, YGL^{+}10, ZCGH11, dKIM^{+}07, AV99, Ada96, AN95, Amb01, ABG96, AR98, BTT90, BE87, BND01, BN91b, BM98, Cha05a, CM12, CV96, CN98, Con00, CS01a, DG04, Der92, EW91, Fro02, GM04a, Gru98, Hae99, HL04, JBJ94, JJP^{+}02, Joh97, JS84, LYMM05, LMT^{+}04, Mad89, Mah85, MMM01, MØ05a, MMMD97, Mit98, Mon96, MD99, MBC^{+}97, MZWZ05, Nor86, Nor87, NK88, Nor90, Paz97, PT05, PT97, PT99, PDG92, RS97a, RMS04, RCT05, Rus00, Sch95, SSRA98, SCM88, SS91, Shy92, Shy93, Smi98, SCKR03, Thi94, Thi95, Tri98, War94a]. technology [WN96, Zan00, DKW99]. Tectual [QCDM06]. Tehran [AS07a, AS12]. TEL [CDS09, DS08, DKW07, KGG^{+}11, NT06, RLKHL12, WKS^{+}10, TCYZ12, HS12a]. Tele [PCH^{+}06]. Tele-Existence [PCH^{+}06]. telecom [Tri98]. Telecommunication [Hor00, HMG03, KMD02, SvS00, AG98, Alb99, CCK95, DOS99, GP98, Hal96, KMN94, PG01, SW97, HS12a]. Telecommunications [SSN08, dSDL04, BB96, Jon00, Jon01, Per05, PVC^{+}99, VDPD96, Del00, She03, dSDL04]. Telematics [AGB^{+}11, DPNM12, HMW^{+}08, LBM^{+}10, MKB^{+}09, OYM07, Ste05]. teleservices [Hut94, Ste94, Hut96]. Teletraffic [BMK08, KHI06, KHS07, MDY07, RMV96]. Television [TOL08]. Telexistence [PCH^{+}06]. TEMA [NSM07]. Tempe [BBS07, SNSR06]. TEMPO [JM78]. Temporal [AH98, Had03, Kam99, MSP^{+}09, MEB05, PZK07, PTM^{+}11, Pup10, Qui04, FO95, FCF92, Hai82, Jen01, Koy92, RH01, BBP89, EJS98, GO94, KSF^{+}03, RH01, BJS99, DFGN12, Jah93]. Temporally [Nok91]. Ten [LL06]. tendencies [BV92]. Tenerife [BHHK08, Sio10, SCD^{+}10]. tenth [Nor90]. Teplá [KK11d]. Term [DHMM96, Hee94, KJ88, RR93, SS89, Baa07, CJ95, Gie05, Gra96, Pfe06]. Termination [SW10, Plu90]. terms [MvOvRdV05]. Terrasini [MMMO96]. Terschelling [TV94]. TES [BS06d, SDH05, BS03a, BCF^{+}02, CGS01]. Test [RCC12, FP90]. TESTCOM [KD05, NBM09, UDF06, GH04, PVTG07, SHUH08, HW03]. Testing [BW05a, BHKR11, BG06, BZU07, BJK^{+}05, BZ08b, CH09, CS12a, Die04, ELS12, GT04, Gol11a, GW06b, HNRW06, HBH08, HM05a, Kul09, MSS07, NZZ11, NMP^{+}04, PU04, SS11a, SS10, UBW06, Yor08, GN05, BCSW10, BF10, GH04, HW03, KD05, NBM09, PVTG07, PSM10, SHUH08, UDF06, WZ11]. Tests [BH08, BJ12b, Dub09, FG10, GW11a, GM07b]. TEX [Des86]. Texas [BBRM98, CW00, DH94, EKT94, GLP99, Goe98, Gup99, HM04, PS97a, PV02, RA00]. Text [Gel04, Gel06, Gel07, Gel08, Gel09, Gel10, Gel11b, Gel11a, Gel12a, Gel12b, PCG^{+}05b, PDK^{+}10, Gel01, Gel02, Gel03, Gel05, HM11a, HR91, MMOS99, Mat01, MM03, MMP05, MM07, MM09b, RG03, SKP00, SKP02, SKP04, SKP06, SHKP08, SHKP10, SHKP12, JM95, LMS96]. Text- [RG03]. Texts [FMPT10, Sta99]. TFM [GO09, DB04]. TFP [PHZ11, PP12b]. TGC [BF08, BS12, DS05b, KN09a, MSB07, WHR10]. Thailand [AAW04, BLJ08, CYCL09, CJ05, CJ06, DA08, FS07b, FNPW05, GGS09, KL95, KKMN07, LLC09a, LLC09b, Mah04, MWK^{+}09, SMEGA10, TKCH09, TNA^{+}10, TKSN11, YT99]. Thalheim [DKS12]. Their [CP10a, CEER96, DKOZ05b, GHSY06, HJ12, JLW07a, Rüt98, YNB08, BJ84, BBP93, BZ83, BRR87b, CER79, DJ90, ENR83, Ehr87, EKR91, GPPW08, HSSY05, KS95, KBD87, Mei89, MMR93, Sal83, Zwi89]. thematic [TAS05]. Themes [CKV08]. Theorem [BF12, KP10a, vEGSW11, JS94, MC94, Pau94, San80b, SWAF95, ST87, AH00, BW03b, BHP95, BNUW09, BDH^{+}99, BJ01b, CMT02, GN98, GF97, HM05b, KK95, MMMO96, MMT08, SB07, SBG04, vGH96]. Theoretic [Des09, Feh11, Kur10, SN08, Smi12, Thi10, ACG^{+}07, HS98c, Bod03, BW00, BKM07, BEFP08, Fom06, GSLM12, HNW04, KK11d, Kuc̆02, PH10, BS92, BL01, GS88, Hof82, Kra05, May93, MST95, Moh91, Moh97, Nag90, Nag95, NDFMS97, TS87, WNE99, vL89, vL94b]. Theoretical [AS02a, AF02, AH03, AKG06, AP07, AKR03, BVNW87, CDV03, CL90, CJ95, CW88, CHK06, DFP^{+}12, DH04, DKN12, DT06b, DKS12, EFW93, EMW94, EN74, EEM^{+}11, FC04, FB12, FR01b, FJ92, FM84, GRS06, Gue90, HMV01, HHL07b, HWLJ08b, KP00, KHG01, LBNF08, LM04, LO10, MP95a, Meh85, MT99a, MVN86, MC89, MMK98, PR03a, Per07a, Per08, Per09a, Per10a, Per11a, Per12a, Pin89, PR96, RRR99, RS06, RT00, RM97b, SG10, Sim92, TW07, AR98, BYGP95, BN91b, CS03, CRW97, CV96, Cho88, DK90, DD05, FC93, FK95a, GPV00, GP89, HF97, JS84, KK87, Len90, LM98c, Mad89, Mah85, Nor86, Nor87, NK88, Nor90, Raj02, RS97a, SKMR94, Sch03c, Shy92, Shy93, Sin94, Thi94, Thi95, vLWH^{+}00, AI97, BBdB12, BCC06, BL03]. Theoretical [CDGW10, CP11, CLP05, CK82, Dam99b, Deu81, FHY08, HM94, HW05, IM91, JLW07b, KSS02, KP01a, LM09, LA05, RRR01, RD12b, TWW77, Wei79, vLWH^{+}00]. Theorie [MOT12]. Theories [But10, BSG03, Cai05, DVVA08, DS06, FBHW00, FMY04, HHL07b, HHL07a, HWLJ08b, HWLJ08a, HZBF10, HZGZ10, HGGG12, HMJG12, JMP12, KPK^{+}10, LOR10, MPV12, MW08, PR07, PBSB09, Qin10, SW08, Som09, SS06e, SPS09, WJ95, BK90, CL01, FBKS03, FRA05a, JL00, MT02b, MWJ97, MSR99, Nie99, Oka03, SRW98, WMT96, ZL98, Zwi89, FCF92, Thi00a]. Theory [AD05, AL11a, AKZ01, Abe10, AK90, AM10, AM91b, AM91a, AI06, ABC^{+}06, ACK^{+}07, ABB10b, AL03b, Ama08a, AC11, ASLM09, AZ11, AF06, AJS00, AL11b, APP12, AL06, AM05, Aus86, AKR^{+}09, BG91a, BH06a, BLS06, BDZ04, BKA11, Bar08a, BH06b, Bar11, BDL08, BDCM04, Bes93, BCI85, BFG^{+}12, Bih03, Bir12, Bod05, BI73, BLvdH10, Boy01, BRT11, BZ09, BJKK02, BG07b, Bue04, BD03, BR07c, CC04, CV07, Can08, Cas09a, CFGN11, Cer05, ČGH^{+}11, CBNR02, CBPS10, CP88, CW09, Cle92, Cle99, CR04b, Cra05, Cra08, Cra12, CVÉ07, Dam91, Dav91, De 07b, DFKM08, De 95, DIMV09, DFMV10, DMV12, Deg03, DV09, DG12, Der04, Des02, DO89a, DO89b, DPA07, DVV07, DN09, DF02, DT06a, Dro08, Due04, DL07, EGMW11, Ehr85a, Ehr85b, EEK^{+}87]. Theory [EKLM87, EL04, EL02b, EM10, Est98, FI08, FK02, Fin99, FBM12, FK12a, FK12b, FW09, FGTZ08, FMRW09, Fum97, GLSY10, GH03, GY04, GL10a, GJT03, GLZZ09, GKB^{+}08, GO11, GK90, Gil10, Gor03, Gor10, GBMS11, GS00b, GHH06b, GP02, GW12, GH07b, Gud08, GQ95, Gün88a, Gup10, GKOT00, GPV88, HP12, HK04, HZ05, HR06, Han98, HMT10b, HKL07, Hed03, Hel94, Hen08, HP06, HHLR08, HPP06, HRSS08, HKLP12, HJW00, Hof11, HM01, HCDL09, Hor02, Hus01, HST07, HSVZ10, ID06, IRM93, IZ98, IZ99, IZ00, Ish11, IKKU08, IT08, Jäh99, JT09, JP04a, Jou09, Kap10, KL88, KS02c, KK11a, KM96, Kim01a, KS02d, KSUZ11, KY07b, Kno11, Knu02, KKS10, Kos98, KU12, KP09, KO07b, KP11, KWT03, KV11, Kur07, KCG09, LC06, Lai03]. Theory [LOX99, Le 02, Lee04, LW11, LP10a, Lin14, LN03b, LLM09, LR05a, LSF07, LS06c, Mai00, Mar11a, Mar11b, MY01, MBH^{+}09, MVOF08, MMP07, Mat09, Mau96, ML11, MP09, Mic10, MMK08, MS08b, MDP04, MDPQA09, MDPQA12a, MDPQA12b, MC98, MHS10, MZ06b, NP02, Nac12, Nao04, Nao07, NP10, NE02, NKM^{+}09, Niv98, O’B12, OiDP98, Oka00, OW07, Ong10, OSC^{+}08, OF15, OST11, PR01, Pal00a, Pat11, PRK12, Per11c, PB08a, Pez03, Pfi01, PSN95, Pie08, PJ12, Pre00, QV90, RR08, Rei09, RSB07, RBsC^{+}06, RT10, RT09, Rov98, Roy05, Rue93, Sag05, Sah13, STP^{+}10, San01, Sas05, SMBQ07, Sch04b, SW03a, SS06a, SZ93, Sei07, Sei12, Ses06, STS04, SBM08, Sma08, Smo00, Som09, SWB06, Ste98, Swi99, Tho99, Tiu00, TZS00, Vad07]. Theory [VPBŠ03, Vau05, Vau06, Wal04b, WS12, Wat00, WK02, WMS04, WTP^{+}06, Wil01a, WS06, WDKK07, WMB10b, YI12, YDKM06, Zhe02, ZT04, dA09, dLZ12, dMS09, vBC08, vHV08, vLIvdH^{+}07, vLMP^{+}10, vdPS08, AH94, AK97, ABK05, AJ94, AI98, AB97, BM99a, Bai05, BSW95b, BB99, BD97a, BF91, BD97b, BR96a, BH92, BD93, Boh75, Bor87, BGR03, Bos00a, Bra75, BWT95, Bra80, BFJV97, B^{+}12b, BBL87, Bud91, Buh98, BR96b, BC98, BE01, BT06, CCD04, CLM03, CC91, CC97a, CFG96, CD05, CP99, CG88, CW89, Coh96, CM05a, CK01, CMT01, CFP94, CDG89, Dam99a, DR05, DD95, DIMV11, DS98, DPA05, DXS91, Dos93, DKN02, Ell95, Esi93, EF03, FBSMS05, FS99b, FC93, FK95a, Fre01, FM99, FM91, GJ93, Gec81, GLM93, GV95, GLM97, GS94]. theory [Hal00, HRR93, HMT99, HW01, HF97, IT03, Jah93, Jan93, JSZ95, JKB96, Jen92, JP94, KSHS01, KL96, Kar77, Kar83, Kil05, KD99, KRS02, KK95, LCS91, LN01a, LT01, LS95, wLN99, Ley01, LM97, Man75, Mar93, MW97, MR97b, MS96c, Mon01, MDBF01, MNS95, NS00, NS81, NU92, Nyb98, OK96, PTB99, PS02b, PN93, Pic86, PMD90, PMD92, PMD94, PMDA96, PMD97, PMDK00, Pit86, PPR87, Pit89, Pit91, PRJ95, PJ97, Rei95, Reu97, Reu99, Reu01, Ric98, Roc92, Roy87, San80a, Sd98, SS94, Sel86, Ser12, SA96, SJT05, Sko85, Ste99, Val94, Van85, Vit95, VBCBS05, WY99, Wil91, WB89, YI91, dB92, ter97, vdBV01, vdAB03, ADDL08, ACL12, ACH00, BW05a, Bau98, BEJvR11, BG06, BMS95, BZ08b, CV11, CCL06, yCCZ07, Can08]. Theory [CC09a, CM09a, CS12a, Cra12, DGR92, EEKR00, GT04, Gre75, HR06, HM05a, HVT12, HdL12, Ish11, IY95, KMPR04, KM08b, Kil05, KLFK10, Kul09, Lin14, MSS11, MSS07, Mic10, MP12, Nao04, OT11, Pai09, Rei09, Sah13, SS11a, SS10, TG10, Vad07, VGP08, ZBRL12, dSOSR04, dSOSR06, vDKS11, Bud85, Had96]. Thera [CD01]. Thesis [Lib07, Gur04]. Thessaloniki [BR07b, BR09, DDI10a, DDI10b, DDI10c, EKP^{+}07, EHH12, MCKB06, MPS06, VS02]. Things [ABK12, FLF^{+}08]. Thinking [MS08a]. Third [AK90, AV99, AJNJ12, ABS01, ASS10, AAA^{+}11, ACP01, AVP07, Akm01, Alb99, AGL^{+}07, AAB^{+}06, APSZ02, ABB99, AFM01, AL03a, AC11, ABS91, ACL98, AB04, AAB^{+}07, AS10, AM00, ABT04, ASAM07, AL06, Ayk09, ALL^{+}10, Bai05, BDK10a, BEOT06, BCC06, BF08, BTQ02, BCM08, BBP^{+}07a, BBP^{+}07b, BS06a, BEP04, BD97a, BP03, BF12, BI03, BFGL10, BL10a, BGPTF10, BBKW03, BB03, BS01a, BHL90, BBZ00, BvDMST11, BMS04b, BHHJ11, BKLR12, BHU03, BKS01, BBR07, BDDS06, Bos01, Bos96, BCK10, BR09, BNZ06, BHV11, BCH^{+}05, BKNS06, Bri97, BtHRBB12, BSH06, BSL^{+}09, BCF^{+}02, BS01c, Buh98, BE01, BL06d, BPS02, BGW06, BKO03, CCL06, CDP02, CK06b, CK06a, CFH^{+}12, CSM10, CMZ99, CMM^{+}09a, CGL^{+}11, CYZ12, CNA08, CD00b, Chi95, CP98b, CP98c, CW99, CGP03]. Third [CPPB06, CCK95, Cle92, CPvdH03, CÓ11, CGS02, CLS07b, CAP08, CEM^{+}06, Cor02, CDFP11, CS96, CH01, Cro03, DU09, DSHL11, DSZ^{+}11, DSM94, DB98a, DW07a, DAB01, DN04, DH97, DW12, DE05, DZ10a, DMW00, DIMV09, DDJ00, DG07b, DHX09, DMLW11a, DMLW11b, DMLW11c, DFG^{+}12, DJNR06, DY10, DS08, DMRS06, Din03, DMB^{+}07, DPP98, DO11, DPSU11, DDS02, DHP09, Dub09, Duf11, DHH^{+}08, ERW99, Ekl94a, Ekl94b, Ell95, EKO^{+}04, EP06, Erl06, EWZ11, EEM^{+}11, EKS00, FV04, FGH98, FS07b, FZJ01, FHFL06, FRSH05, Fra99, FGL96, FCLG04, FS06, GP07, GVCA07, GW96, Gel02, GB97, GFV09, GOW03, GMI06, GP09b, Gol94, GKS05, GLM93, GN08, GW12, GO06, GAR04, GB07, HNH^{+}09, HR06, HMT10a, Ham91, HKB99, HLR^{+}98, HS06, Har01, HH10, HSS07]. Third [HRTR05, HRSS08, HMP99, HJRS99, Hol07, HRRY08, HT06b, HPZ10, HSSV12, HHL07b, HHL07a, HS99, HS08c, HJ00, HS98d, HKM^{+}11, Iba92, IZ00, IFGH10, IKY05, JRL09, JK00, JNH11a, JNH11b, JPXS04a, JP06, KWA01, KCYL08, KL07a, KJ09, KV04, KM08b, KP04, Ker01, KL12, hKAS^{+}11, KR00a, KSW99, KNP01, KMSS02, KKS10, KMVY06, Kro00, KF12, KH09, KZ10a, KLT09, KCK08, LV12, LdRV94, LSW01, LOR10, wLN99, LKK^{+}07, LGG97, Leo04, LNE06, LHDD08, LPH00, LDXZ10a, LDXZ10b, LMC12, LZ05, LM98c, Luo06, LK00, MJYT06, MOH^{+}12, MCK^{+}06, Mai00, MG08, MBS11, MZ07b, MWY01, MR01, MVC07, MBH^{+}09, MNW01, MBMS07a, MBMS07b, MTCOBH11, MTO99, MF08, MPR07, MG07a, MMM10, MPvH04, MS02b, MSW02, MD96, Mil00, MFÁ^{+}09a, MFÁ^{+}09b, MKR^{+}04, MS08a, MAFSS04, Moo01]. Third [Mor95, MBH^{+}11, NP12a, NLKJ06, NKJ11a, NKJ11b, NW97, NCBT04, NBK04, NABK12, Nor04, NQR09, NU92, OKHH^{+}08, OA93, OSRS07, OZ09, PDH99, PZE^{+}08, PDSN12, Par02, PCA^{+}09, PdRM^{+}12, PYLW08, PRT06, PD04, PR03b, PS08a, PCS^{+}07, PBGK03, PU04, PTZ03, PY05b, Pet92, Pfi00, PMD94, POS00, PZ09, PG01, PK12, PB08b, PLG06, Pol00, PS09b, PMS08, Qin10, QOZ01, QS00, Ram01, RS09a, RUL04, RM02, RAW07, RTZ08, Rau04b, RBF06, RM04a, RGW01, RLT^{+}08, RK02, SP09, SN08, SdSMO12, SMB03, STT11, ST08, SH12b, SF74, SKP10, SR07, SNZ08, STT04, Sen10, SG09, SW09b, She03, Shu09, SU97, SBD03, SL99a, SF96, Smo97, SKP00, Spi95, SDK^{+}08, SHS10, SFV^{+}11, Stü09, Sub05, SV01, SHO99, SKN05, TMdB11, TSG00, TSJ12a, TSJ12b]. Third [TN09, Tiu00, TO08, Tor93, TNVDF06, TG10, Tra04b, TMN09, Tro11, TCLH10, VJKT00, VDPD96, VU09, VBS^{+}10, VS05, VSTN06, VP04, WHL09, WYZ06a, WYZ^{+}06b, WJS^{+}06, WLGB^{+}08, War94a, Waś96, WPM07, Won01, WYX08, Yak04, YJMU06, YJS^{+}06, YAF^{+}02, YM01, Yor08, YSL12, YKM06, ZCCL10, ZS00, ZOCJ07, ZZ99, ZZTG07, ŻR99, dMS06, vDLJ11, vdBK05, vdMBD08, vM94, ABLM05, ADB^{+}05, Apo92, BUG93, BCT06, Bod96, BBD97, BCKT05, BV94, CTT93, CCH02, CAZ05, Cou93b, CKL05, Deh93, DF96, Dos93, EFM04, EKL^{+}05, FS05, Fer96, FLMS05, FvdHJT10, GG05b, GWS05, GS05, Gol96, GS94, HSSY05, HIS05, JL97, KL92, Knu99b, LM93, LVW05, LSL04, LKZCS05, Mal95, MNT95, MW03, Mol95, MBA05, NM94, PCG^{+}05a, PFCS05, PT05, PV04, Ran93, RKM06, Ste95, SSKR02, X^{+}12a]. third [Yi05, dBBGdR05, Bra73, Sha86, SSAP05a, SSAP05b]. Thirteenth [CCR86]. Thiruvananthapuram [CM12]. Thomas [CGMQ08, KSW03]. Thought [LLM09]. Thyrrhenian [Pal01]. Tianjin [ACK^{+}07, HNS^{+}12]. Tiburon [Ric02]. TIC [Har01]. TIC’98 [LO98]. TICTTL [BvDMST11]. TIDSE [GMI06, GSH^{+}04]. tight [Bol99]. tightly [Nag96]. Tilburg [BB01b, vdHP12]. Time [AH98, BR97, BCDN07, Cha85, CKH10, Dad01, DO02, Die04, DFGN12, Dwa00, GKL^{+}11, JLW07a, Jos00, LY04, O’H98, PA01, SKP07, VEG^{+}11, AW05, BC04a, Had94, Hoo91, JM78, JP96, Jos88, Kat99, Koy92, LdRV94, Mal97a, RR98a, Vyt91, dBdRR89, CH04, Haj96, KPR07, MP10a, Lis89, dB92]. time-critical [Koy92]. Time-Series [DFGN12]. Timed [AB06, CJ08a, CH10, FT11, JN12, LN04a, OV09, PY05b, RT07, LY04, Rus05]. Times [DS08]. Timetabling [BD03, BR07c, BR96b, BC98, BE01, BT06]. Timing [AGCP^{+}11, AS07c, HAB02b, MPK04, MvL10, SPB00, SM09, VAM06, vLS11, CM03b, PVV05]. Tiruchirappalli [KA12]. tissue [AD03]. Tissues [LST^{+}12]. Title [PA03b, ZYH03]. TLCA [Cur09, Ong11, Roc07, Urz05, Abr01, BG93b, DH97, DCP95, Hof03b]. TLCA’99 [Gir99]. TMA [DPSU11, POP09, PSD12, RMB10]. TMN [Hal96]. TMN-based [Hal96]. TMRA [MP06, MSG07, MG08]. TN [DGSM11a, DGSM11b, DCFR12, GY08, Gie08]. Tobago [DD07]. Today [MYS10, WV99, vL95]. Tohoku [NS81]. Tokyo [ANS04a, ANS04b, ANS04c, AKU01, AK03, AKT06, AF99b, Asa99, Bar03, BIRT05, BGY96, CHM^{+}10, CQSS09, DK02a, DK02b, DSL08, Dos93, FTF88, FTF89, FTF91, FMY04, Gel11b, Gel11a, Gom01, ICSY07, IS02, IN11, Jan93, JSZ95, KUA00, KM08b, KU01, KL02, Liu01, Mat09, MNTY05, NMM10, NYHO07, NMPM08, OT11, Oka03, OS07, OBM11, OKdQ09, PLI08, Sah13, SKS^{+}06, Sak91, SFS11, SKSA12, SILS05, SHYEF08, SHUH08, TOOO07, TO08, VJKT00, VJAA03, Wad86, Wad87, WHL09, WSTI07, WY99, YKKZ07a, YKKZ07b, YK82, Yas00]. Tokyo-Odaiba [VJAA03]. Toledo [Her00, NMP^{+}04]. Tolerance [BKNS06, MCER10, Sch12, BL94, BHS05, BJRT09]. Tolerant [BJRT06, DO02, Jos00, JP96, Jos88, LY04, LdRV94, ON80, RR98a, Vyt91, SS90]. Tom [JPR04]. Tønsberg [DFL^{+}10]. Tool [PF07b]. TOOLS [BV11, FN12b, Vit10, BWAM09, BB95, Ber12, BK04a, BJRT09, CT02, CG04, Cha05b, CCAS12, DSLT08, DSLT10, DSHL11, DMB^{+}07, EV01, GNTF02, HM11c, JMP12, KKM^{+}11, LOR10, LS05, Mar11a, Mar11b, MW08, MdPA98, MSA99a, MC98, PP12a, SW08, SBM08, SMEGA10, SS07c, Vos03, Wal84, Wes99, WCBD08, ASP85, AAN00, BF91, Bre03, BJ95, DM01a, DHM00, DC00, Fie02, Gac02, GK01, GH93, HK94, HBS00, JHC02, KS03c, KMTU03, MR97a, MB98, PSS98, RD84, WB89, Fie02, HBS00, KS03c, AL11a, BvDMST11, Bri95, Bri97, Cle99, EM10, FK12a, GH03, GS00b, GH07b, HZ05, HP06, JP04a, KS02c, KP09, MS96a, MY01, MN12b, RR08, Ste98, Urs02]. Top [Sch11]. Topic [MP06, MSG07, MG08]. Topics [BCC07, CUdlC04, DKRT06, Kut12a, Kut12b, MOBS06, MMG10, AP93, AS98, FG96, Foo99, LM98b, MST92, MS00b, Nos88, Sat97, Sch92, Abe06, Dun12, ETG02, Fis09, GS95, Joy03, Kia11, Mal08, Men05, Nac01, NP12a, Neu04, Oka04, Pie10, Poi06, Pre02, SF08a, WAR94b]. Topological [Ren02]. Topology [FFL^{+}12, Ehr89]. Torino [BC96a, BCD04, BDd09, CM03b, DSLT10, DCC^{+}11, DBD^{+}11, LM03, MH11, PRG11, RRR01, SFN^{+}11, Tor93]. Toronto [ASR^{+}07, AMRŚ08, ALPH10, AHSC11, BMR^{+}04, BGJ11, BC98, BHM^{+}02, BTF05, CCS04, CKP^{+}12, CBG^{+}10, CDN12, DFPV11, Dwy01, FMP11a, FMP11b, FMP11c, Fri04, GK01, GDG11, Jac04, JP04b, KC05, KI12, LLT07, LMCH12, LN01b, LSIT11, MDHB11, MS05, MRP11, McC96, MV12a, MGSM12, NP12b, NOOS05, OSV11, PMWO02, RS12, SDH05, SW08, SWSY11, VY01, Vos03, Wal04a, WvdH03, WG11, YLW^{+}07, YSP^{+}10, YSL12, vBC08]. Torre [COBV11, CDFP11]. Torremolinos [BHV11, CRJ11a, CRJ11b, FN12a, HC11, MS95]. Torremolinos-Málaga [BHV11, CRJ11a, CRJ11b, HC11]. Torun [MS08a, WDKW12a, WDKW12b, OT08]. TOSCA [MP12]. Touch [KMSS02]. Toulon [CW89]. Toulouse [ABD^{+}99, AFGvE11, ALR11, Aus93, AB97, BST03, BB01a, BKLP11, BL89, BTF02, BBMSD09, Cha09, Cup00, CD11, COB^{+}08, Dam99b, DMW00, DH10a, DEQ92, DOS99, EH08, Est99, FL00, FKP99, FHP03, FLP11, GLV09, GOZ05, HT97, HLSZ11a, HLSZ11b, HT11, KT11, Mal02, ML04, MVL11, MM97, MPS04, NL99, OO12, PAD^{+}08a, PR84, Pol86, RS03, SA97, Wan98, WJP07, WFB12, WWD99a, WWD99b, dCHM12]. Tourism [CFVCF^{+}12, FdCLN^{+}11]. Tours [CM03a, MKB^{+}09, MTC^{+}08]. Town [ABGL09, KJR05]. Townsville [McC97, PGD07]. Toyama [TM09]. Tozeur [LMM10]. TPC [NP09, NP11a, NP12a]. TPCD [KK95]. TPCTC [NP09, NP11a, NP12a]. TPDL [GBMS11, ZBRL12]. TPHOLS [SBG04, BNUW09, HM05b, MMT08, SB07, AH00, BJ01b, BW03b, BDH^{+}99, CMT02, GN98, GF97, vGH96]. TPNC [HVT12]. TPPP [IY95]. TQC [CM09a, KM08b, vDKS11]. Trace [Van85]. traces [Die90]. Track [Abe06, Dun12, Fis09, Joy03, Kia11, Mal08, Nac01, Oka04, Pie10, Poi06, Pre02, GHK97, Men05]. Tracking [FK09]. tracks [ZVU99]. Tractable [Cad95]. Trade [OM98]. Trading [LSJ06, NS99]. Traffic [MDY07, SHS10, VS09, DPSU11, POP09, PSD12, RMB10, VS09]. Training [ECV^{+}10, EEM^{+}11, GLW07, GMUW12, IJ06]. Trakhtenbrot [ADR08]. Trang [TS10]. Trani [FPP11]. Trans [Fre98]. Transaction [CZZ08, GHR89]. Transactional [Nag01]. Transactions [PSGB^{+}04, PS05a, RA06a, RA07a, Fre98, GT08, GTW^{+}08, GT09a, GTM09a, GTWC09, GT09b, GT10, GTA10, GTM10a, GTM10b, GT11a, GTSS11, GT11b, GTM11, GTP12, GT12, HKW09, HKW^{+}10, HKW11a, HKW^{+}11b, HKW12a, HKW^{+}12b, JvdAB08, JvdA09, JBK09, JDK10, JDK12, JvdAM^{+}12, KOFJ09, KMK10, KMSJ11, LCH^{+}12, NK10a, NK10b, Ngu11a, Ngu11b, Ngu11c, Ngu12a, Ngu12b, Ngu12c, NJ09, NJA^{+}11, PCME08, PCME09, PCMC09, PCM^{+}10, PCMY11, PCM11, PCM^{+}12a, PCM^{+}12b, PS05b, PSD^{+}05, PS06, PSD^{+}07, PSM^{+}07, PS08b, PSR08, PSW^{+}09, PS10, PSS^{+}10b, PSC^{+}11, PSS^{+}11, PS12b, PMGO05, Pri05, PZ05, PCE06, PHPL06, PP06, PIMN06, Pri07, Pri08, PDAN08, PBP09, PBG^{+}10, PBPdV11, PPdV12, RA06b, RA07b, RO09, SST00, Shi06, Shi07, Shi09, Shi10, SEK^{+}11, Shi12, SK12, Ste07a, Ste09a, Ste11a]. Transactions [Ste11b]. Transfer [ABC^{+}06, CUdlC04]. Transformation [Alp11, BEH^{+}85, Bru04, CEKR02, De 10b, Han09, Hil06, Kin08, MSS02, Pet02, Pue07, Vid12, Bos00b, Eta05, Fle99, FT94, Fuc98, Gal97a, HKB93, Lau01, Pro96, BR97]. Transformation-based [BR97]. Transformational [FLVS11, LSV06, LVS08, Yu02]. Transformations [CV11, CEM^{+}06, EHRT08, ERRS10, EEKR12, ELS^{+}10, HdL12, LS05, Pai09, PNB04, Rüt98, SNZ08, TG10, VGP08, EEKR00, Ehr04, NSM00, SE94, SVV12]. Transition [Geu98, Pup10]. Translating [Fre87]. Translation [FGH98, FT04b, Whi00, Bra76, Ric02, Ste95, Amt99, Yel88]. transport [Hut94]. Transportation [AMZ09, Leu03]. traveling [Rei94]. treatment [JM78]. TREC’98 [LM98a]. TreDS [SLPM96]. Trees [AM91b, Arn90, DN88, DO89a, Ehr85b, EEK^{+}87, FZ86, Kir96, Rao92, AB81, AP83, Tis94]. Treewidth [Klo94]. Trends [APSZ02, AIL12, BSW95b, BEJvR11, BFG^{+}12, BBK^{+}11, BHK01, Cai08, CB11, CR01a, ČGH^{+}11, CGBZ08, CS10, CM09c, DF10, DBSvD09, Dig11, FS07c, GKB^{+}08, GHH^{+}06a, GP02, GHI^{+}06, HK12, HJW00, HZ08, HSTU99, IS02, Jac09c, JKB96, MK12, Mue06, Nie09, NKM^{+}09, OA07, PR01, PTB99, PJ97, Rov98, Shu11a, SAUL11, SKR^{+}10, TSKGB04, VPBŠ03, WTP^{+}06, WPH03, WV99, YW06, YYS^{+}12, ZGY04, ZO10, dMS06, vLIvdH^{+}07, vLMP^{+}10, Apt00, BCM00, BC93, Ehr91, EO94, Fia99, FMO05, HOD96, Kum90, LMT^{+}04, Mau91, MCP93, Pal01, Par98, PS97b, Pod95, PS98b, RTA95, SKMR94, ST88, VBCBS05, ZY01, dBRdR86, vL95, AZ96, AGS91, BJSW09, DW06, DW07b, DL81, DRSS01, GPHF^{+}10a, GPHF^{+}10b, GPHF^{+}10c, HSTU99, Jan88, KK87, Kut12a, Kut12b, LM98a, LPH00]. Trends [PHZ11, PP12b, SUW02, SLPM96, ABGP05]. Trento [Bat01, BCG^{+}11, BP01, BAL88, BSB^{+}99, BSS00, CS12a, Dal92, Dam99b, Gan99, GS09a, HP99a, HMPZ09, MST12, MBW08, NR99, OWPY03, PH12, Pri06, SBV^{+}11, SS11b, Stü09]. Trevisan [Gol11b]. Triangulation [GTM11]. Tricks [OM98]. Trier [ABMH08, DFMV10, MT99a, TG12]. Triggered [RG03]. Trinidad [DD07]. Trinity [FHM82]. Trois [ELN^{+}10]. Trois-Rivières [ELN^{+}10]. Trondheim [AK10, ABS91, AB03, BK05b, BCH^{+}05, BS06d, Col07, CDW86, Der92, Din04, DW06, JP05b, KS04b, KZ97, KOS07, MMK08, MBJ^{+}07, PGBK04, Pie05, RK93, SPH07, THT03]. Troy [MMM10]. TRUST [ASS10, DsSL^{+}07, FBSMS08, KWC^{+}12, MBP^{+}11, ZSY^{+}06, APP12, BDE12, BFMS08, BFJM09, CMM09b, CLBP11, DGM09, DG10, DEG11, DG12, DMRS06, DMRS07, LSK08, MFG12, ASS10, FST01, FBKS03, FBSMS08, FHFL06, FHLP09, FHKQ12, FKL08, FLP11, HIS05, JPD04, KLP04, KLS10, KWC^{+}12, LPT07, LPK05, MBP^{+}11, NT03, SWMM06]. TrustBus [FHFL06, FHLP09, FHKQ12, FKL08, FLP11, KLS10, LPT07, LPK05, KLP04]. Trusted [CYM^{+}11, LSK08, NRWI09, RJS^{+}08, XYM^{+}07, XBS^{+}10, YJMU06, CMM09b, CY10, CY11a, CYZ12, LSK08]. trusting [FBSMS05, FBSMS05]. Trustworthly [WHR10]. Trustworthy [ASS10, CJR^{+}10, KWC^{+}12, MBP^{+}11, RSS06, BF08, BS12, DS05b, KN09a, MSB07, WGMP09]. Truth [MR91]. Tsakhkadzor [GKMV10]. TSD [HM11a, MMP05, MM07, MM09b, SKP06, SHKP08, SHKP10, SHKP12, Mat01, MM03, SKP00, SKP02, SKP04]. TSD’99 [MMOS99]. TSDM [RH01]. TSP [Rei94]. TSPack [Cha85]. Tsukuba [HML97, KILW10a, KILW10b, MMT97, MKT98, MS99b, Nip98, NABK12, SHM^{+}10, TNM05, Yi05, YMY^{+}10, vdHI99]. TTIA [CUdlC04]. Tubingen [SH91, Sch92, CC09b, GBC^{+}09, PRG09, VGM^{+}09, BLPW02, RBGS04]. Tucson [Apo92, Bat09, Che03a, CMZL04, CAB10, EOR06, Kru08, RBsC^{+}06]. Tuebingen [BvLR04]. TUG [SBH^{+}04]. Tulane [Mai88, Mai90]. Tunis [BCC06, YNB08]. Tunisia [BCC06, DYCD09, LMM10, Mel07, Per10b, Pre09, YNB08]. Turin [AL06, BDG12a, BKW08, DD95, DCM82, FKL08, Ham08, Mor06, PAV^{+}10, PW08, SEN08, TBE08, WSF08, vdHCD07]. Turing [CDL12, Sze94]. Turkey [ADK04, BdBK03, CM10, DCC^{+}10, DBD^{+}10a, DGR06, DFS08, EAES^{+}10, FHY08, GJTS06, HWW^{+}10, HH10, JOD98, JP12a, Kay03, KMS12, LSY^{+}06, PSE10, PRG10, RBGG06, SMD04, SGSV10, Sav06, Sma08, ÜÇA10, Yak00, Yak02, Yak04, YN06, Yal10, YŞ03, YGGÖ05]. Turkish [Sav06]. Turku [BBKW03, BPS02, CKPR11, CMS08b, DKLS04, DT06a, HKL07, NR04, PS02b, SGPP06, SS77, vGH96]. Tutorial [AGM05, AG07, ABG09b, AG11, ADMH04, ADM^{+}06, ADM^{+}08, AAB^{+}07, ABW10, BBF^{+}06, BBM^{+}08, BC08, BBPP09, CT02, CG04, DF09, FG01, LMŠT01, MN12a, MN12b, PdA^{+}11, TFE^{+}09, DN93, FG04a, GG98, JM95, LMS96, Pit86, Bra88, GH92]. Tutorials [HK12, PT91b, GAB02, dBRdR86]. Tutoring [AKM10a, AKM10b, CGP02, CCPP12, IAC06, LVP04, WANL08, FGM92, FGL96, GFV00, Goe98]. TV [CCJ07]. Tver [NT92]. Twente [AAR05]. two [Wag94]. TX [BBA12, BMSW12, Cha05b, CMCS11, DG06b, FD11, JS11a, KS09a, LHDD08, RL11b, ACK00, Gac02, HJ00, Jay04, Mau89, MO92b, Rau04a, SV01]. Tyne [CK01, DMRS06, GH06, HS08a, KU12]. type [BC93, Ehr91, EO94, HOD96, RTA95, ST88, She85]. Typed [BG93b, DH97, DCP95, DL95a, KO91, Abr01, BG93b, Cur09, DH97, DCP95, Gir99, Hof03b, Ong11, Roc07, Urz05]. TYPES [AM07, BDd09, FPMW06, MSH08, BCD04, BC93, CDNS00, Ehr91, EO94, HOD96, RTA95, ST88, BHR90, KMP84, Sch97b, ANR99, AM07, BN94, BC96a, BCD04, BDd09, CLMP02, DNS95, FPMW06, GW03, GPM98, Har01, LO98, MSH08, She85, US90]. TYPES’99 [CDNS00]. Typography [SBH^{+}04, HAB98].

UAHCI

[Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c]. Ube [Nak92]. UbiComp [BIRT05, DF06, KASS07, BIRT05, BH02, DMS04, DSM03, DF06, KASS07, ABS01]. Ubiquitous [AHSC11, ACHH12, BH02, BGR08, BMC08, CDM04, DMS04, DsSL^{+}07, DSM03, DF06, EYX^{+}05, Jac09a, KGG^{+}11, KASS07, KBC^{+}09, KSG^{+}07, LN09, MBH^{+}09, MPPU11, ONPR07, SMBQ07, SHX^{+}06, Ste09d, TG00, VFC08, YGGJ04, ZSY^{+}06, ABS01, BIRT05, Gel99, Tri98, YAL^{+}05, BDGM05, HYMZ11, ICSY07, IMY^{+}07, MJYT06, MS10d, MNTY05, SZR^{+}08, SKCK07, YKM06, YLC^{+}10, ZPTI09]. UC [ACD^{+}07, CDP^{+}05, CDP^{+}06, CCF^{+}08, CCD^{+}09, CKPR11]. UCHPC [ADB^{+}12a, ADB^{+}12b, GVT^{+}11]. UCMA [hKA10]. UCNC [DLJ12]. UCS [ICSY07, MNTY05, YKM06]. Udine [AB78, BBD^{+}09, CC91, Chi03, CDM04, KLT09, dlBP08]. UDM [YMY^{+}10]. Ugo [DDM08]. UI [Ayk07a, Ayk07b]. UI-HCII [Ayk07a, Ayk07b]. UIC [HYMZ11, IMY^{+}07, MJYT06, SZR^{+}08, YLC^{+}10, ZPTI09]. UISW [EYX^{+}05]. UK [AJS05, AM91b, AM91a, ABFK10, AvAS06, AvASD06a, AvASD06b, AvASD06c, AF11, AM07, And96, ATO^{+}09, APHA11, ABM05, APP12, ABRZ06, Atl08, ACR^{+}08, ALL^{+}10, AKR^{+}09, BBMB^{+}08, BW05a, BGHJ07, BBP89, BMPM09, BMH05, BBLT06, BH06c, BE08, BR06, BGF06, BW10b, Bla05, Bod05, BDS04, BBBB08, BF06, BF10, Bra09a, BW01, BMS01, BEFP08, CG12, CDP^{+}06, CFB^{+}05, CMSM^{+}10, Cas09a, Cer05, CZ05, CW09, CRW08, CMM09b, Che11, CCMR05, CCMR06, CCMR07, CCMR09, CMMR11, CCMS11, CMSA12, CPPB06, CK07, CDL12, CFP^{+}09, CKC11, CPR12, CR05b, CFS06, CMPW12a, CMPW12b, DG06a, DSLT08, DMS04, DJAP07, DG07a, DC12, DS05b, DV09, DJNR06, DMRS06, DP12, DS06, EL04, Erl06, ER05, FGB11, FHRR05, FvdHKL06, Fis08, FHT05, FDC12a, FDC12b, FG06]. UK [Fog94, Fog96, FKBP11, FTC^{+}10, GHJ06a, GHJ06b, GHJ06c, Gal07, GP08a, GGK^{+}06c, GGK^{+}06b, GGK^{+}06d, GGK^{+}06e, Gib12, GH10, GH06, Gol96, GMS12, GJS08, GW06b, GS08, GAM^{+}11, GAN05, GHK^{+}10, HZ05, HMS09, HS05a, HRC06, HS08a, HMTH10, HL04, HSP^{+}12, Hil06, HSKB09, HdQ08, HRM^{+}10, HM05b, HK07, IS11a, JNS05a, JSO11, JN12, KSG^{+}09, KG12, KSS08, KP10a, KLT^{+}12, KRP06, KKM^{+}11, KW07c, KU12, KP09, KP11, Kul09, LSJ06, LMR^{+}06, LJR^{+}10, LB05, LOR10, LNE06, LLLL08, LS06a, LYCV08, LNS11, Liu11, LST^{+}12, Mac12, MOP^{+}08, MBS05a, MSW07, MB06, MHK05, MS10e, MR11, MTNL12, Myc95, ONC^{+}11, Ong05, Par09, PMH09, PCG^{+}05b, PPR87, PJ12, PZH96, PPH07, QQ11, QCDM06, RW11, RS05, RMS04, RV11, RBSL06, Sag05, Sas05, SKP10, Sch06, SC11]. UK [SJHJ10a, SJHJ10b, SJHJ10c, SJHJ10d, Sex09, SCL^{+}06, SPON08, SSAP05a, SSAP05b, Sma05, SMF^{+}11, SMCM07, SG07, SS10, Tah09, TTM10, Tho11, TT06b, TCJ10, TKHS05, TMN09, TMP04, VW05, Wal99a, WT07, WR07, WMdLA10, WNL05, WCBD08, YHR^{+}09a, YHR^{+}09b, YBL^{+}04, YTC^{+}07, YWRS11, ZS05b, dA09, dBM10a, dBM10b, dMS09, dMGFS11, vDKS11, Mez11, ARS03, Akm01, AS03, AD98, AFS02, And94, ACH00, AFL03, AT02c, Bac00, BCG02, Bak95, BMP00, BEP04, BND01, BMW93, BF00, BMS04b, BBH01, BJ01b, BHT97, BDGK00, Boy95, BM03c, Bra02, BD04, BR96b, BC96b, BC96c, Cag00, CLMP02, Chr99, Chr00a, Chr01, CCMR02, CCMR03, CHA03b, CHD98, CH91, CK01, CM00, CS97, CP02, CGvR02, CSSW04, CH01, CF00b, Dar97, DFR02, Dav91, Dyc94, Dyc00, ENP02, Emb98, EW04]. UK [EKS00, FRS02, Fie02, Fit93, Fle99, Fog95, Fri94, GY04, GW96, GM89, GHK97, GL03, GH04, Gru98, GW93, HDC99, HP99b, HV03, HAB02a, HHM97, HB97, Hon01, HP99c, IKR00, JLY03, Jen92, JPD04, JJ03, Joh01, Jon00, Jos88, KMT00, Kit88, KR01, KN03, Koz04, Lau01, LSE02, LJ00, LCB97, Llo00, Lou92, LCG97, LIH99, MPP94, MM01, MA03a, MT04b, MM04, MSS04, OPP04, OWM04, PHJ03b, PHJ03a, PC99, Pat03, Pit86, Pit89, Pol00, PR98b, Rai03, Rea01, RSL01, RSK03, Sak01, San94, SBC^{+}94, She01, She03, SLL01, TC99, TN03, TG00, TBH03, Tis94, VZ99, WS02, Wat95, WM04, WM03, WR03, YEY04, YAF^{+}02, YI91, vdBV01]. UK/Japan [YI91]. Ukkonen [EMO10]. UKMAS [DLFP02]. Ukraine [Pfe94]. Ulm [BFP04, DEKR09, Kar91, SM06, ST12b]. Ulsan [HJL^{+}09a, HJL^{+}09b]. UMAP [AK12, DKC10, HMPZ09, KCMO11, MMDN12]. UMC [CDP02]. Umeå [Kåg98, KEDW07]. UMICS [BDGM05]. UML [ALS^{+}05, DMB^{+}11, HRK^{+}07, HP09, JSdSA05, RR01, RBsC^{+}06, SPC^{+}08, TDK^{+}10, BM99b, FR99b, GK01, JSdSA05, JHC02]. UML^{≫} [BSMM04, SWB03, EKS00]. Uncertain [Heu04]. Uncertainty [BG11a, BMYZ91, BMYZ95, DH10a, GP09b, GL08, HKH10, HLFS12, Liu11, Mel07, PS07b, QCDM06, SC09b, THL11, VY93, YBM87, BB01a, CKPB07, CKM93, DvV96, FK95b, God05b, Had94, HP98, HP99c, KS91, MP94, NZ03, WP90, BMYZ91, SYBM88, YBM87, dCdF^{+}08]. unconstrained [Naz94]. Unconventional [ACD^{+}07, BFGM05, CDP02, CDP^{+}05, CDP^{+}06, CCF^{+}08, CCD^{+}09, CHM^{+}10, CKPR11, DLJ12]. Underlying [MÁ05b]. Understanding [DGSN11, ERK07, Fen00, Mit06, NM08, SGSV10, SL11, SdSMO12, SÇS12, HR91, Hir81, Kom97, Hel08]. UNESCO [DFS08, DFM11, DFT12, MSZ90]. UNICORE [BFT^{+}08, CAS^{+}09b, LMSS07, LAF^{+}10]. unification [BK90]. Unified [BSMM04, EKS00, JHC02, Koj91, SWB03, BM99b, FR99b, GK01, JM78, Naz94, Tha84]. uniform [Rus05]. Unifying [PBCF08, But10, DS06, EJPR01, LBO05, Qin10]. United [AM03, BGM94, Bow94, CC95, DCP95, EHRT08, GK95, Gol94, JBJ94, Lom97, Lou94, ML95, MK96, PRJ95, RT95, SSO^{+}11, SOS96, Sha86, Sma97, Wat95]. units [Ful75]. UNITY [Rao95]. Universal [BMC08, Sch03b, SS04, Ste07d, Ste07b, Ste07c, BMP90, Lev92, LPH00, CS03, FCLG04, Sch98c, Ste05, SP07, Ste07d, Ste07b, Ste07c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c]. Universality [DLM07, MR01, Mar05a]. Universe [MW03, WW02, ZEL04]. Universität [BHR84]. University [AM03, BRW85, Fur95, FU96, GM89, KT96, KC84, Lau93a, Loe74, Mai88, Mai90, NS81, Pal00a, Pat90, PC99, SS77, Sel86, Van89]. UNIX [BGW93]. Unobservability [Fed01]. Unobstructed [Akm87]. Unreliable [KS07b]. Unterricht [BHH74]. UNU [AM03]. UNU/IIST [AM03]. Upon [CK01, DMRS06, GH06, HS08a, KU12]. UPP [BFGM05]. Uppsala [Asp98, BNS00, Bru04, EDGP11, GGK^{+}09a, GGK^{+}09b, HNH^{+}09, HCF06, JP94, JP96, KSW99, PY05b, Pet92, TSKGB04]. UPWN [M^{+}06, THX^{+}07]. Urbana [AGP08, BC09a, BHL^{+}03, KS03c, MR08]. Urbana-Champaign [BHL^{+}03]. Urdiales [IT06, MSX10]. Urgench [EK81a]. URSW [dCdF^{+}08]. Uruguay [BV05a, BPS11, BBPP09, GPV00]. US/Japan [HI93, IH90]. USA [AD05, Abe06, AGP08, ABF^{+}07, ASE^{+}07, ASP01, AKM10a, AKM10b, ADJ^{+}01a, ADJ^{+}01b, ABB^{+}10a, ANS^{+}09a, ANS^{+}09b, ACW07, AH96, AP10, APN08, AN07, AN09, ASAM07, ACHH12, ABRS06, Ayk09, BBMB^{+}08, BBD^{+}10, BC09a, BL05a, BHL^{+}03, BSvRW08, BJ06, BBS07, BKA11, BPCS06, BA07, BBH^{+}10, BBA12, BGP08, Bat09, BBKP06, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BdQ11, BC10a, BGKY08, BG11a, BF12, BKH^{+}09, Bes07, BG06, BMSW12, BRST08, BHHJ11, BR12b, BC07, BFMS08, BFJM09, BGPR10, BRT11, BB11, BGHS05, BAdVR08, BBH08, BKM06, BKM08b, BG07b, BD08, BT06, BTMI11, BHLP07, CJ11, CCJS07, Can08, Can12, CPFD^{+}03a, CPFD^{+}03b, CS11, CS05a, CR05a, CC07, CSM10, CGBZ08, Cha05b, CGKM11, CJRR07, CJRT05]. USA [CS05b, CD05, CU08, CGS07, CBT07, CAB10, CM05a, Con97, CMCS11, CMB^{+}08, CDA^{+}06, CEER96, CFB11, DGSM11a, DGSM11b, DL06b, DSRW08, DG06b, DLT05, DCFR12, DKC10, DGJ^{+}05, DHRT10, DS09b, DPB^{+}04b, DIS11, DG07b, DJNR09, DL04b, DJ08, DCFY10, DDI09, DLW06, DF06, Dov05, Duf09, Duf11, DG05b, Dun12, Dwo06, EJM10, EK09, EFM04, EGMW11, EM08, ELM05, EdS08, EOR06, EN06, EOSW07, EG10, EW05, EIRW06, ET06, FV04, FV05, FRA05a, Fei03, FRS05, FFRS05, Fer96, FW07, Fis09, For88, FILP06, FS08, FS10, Fra04, FKM08, Fre98, FHP11, FD11, FK09, FS06, GVO^{+}10, GKL09, GPCL10, GY08, GRB^{+}07, Gho10, GH07a, GAA06, Gib89, Gie08, GSYZ10, GS09b, GH12a, God05a, GP09b, GJRR08, GJN10, GZ09, GA09, GJRR11, GGHS07, GPPW08, GCM08, GP12]. USA [GQ11, GHP09, GYA^{+}06, GM11b, GF12b, GW11b, GM08b, GJRS12, HMRT05, HR06, Hal09, Har09, Har11a, HNRW06, HMP08, HCS^{+}05, HMW^{+}08, Hen06, HC05, HT06a, HW09, HRTR05, HRR^{+}06, HSB^{+}10, HW06b, HM04, HKM^{+}07, HMBD09, HW08b, HKS09, HMT10c, ID06, IR08, IB07, IJ06, IKY05, Ish11, Jac09c, Jac09d, Jac09a, Jac09b, Jac11a, Jac11b, Jac11c, Jac11d, JSV11, JW05, JSS08, JT09, JS11a, JMO09, JP11b, JS02, JDMB10, JMP12, JP12b, KS09a, KJR05, KMR^{+}05, KL07a, Kar09, KS12b, Kia11, Kil05, KS06, KS07b, KSHM09, KPR07, KP15, Kru08, KS08b, KS10b, KR12, Kur09, Kur11, LP97, LSgL^{+}06, LN09, LOTY05, LPH09, LHDD08, Li11b, LMT00, Lin14, LKT08, LBDW09, dVLKR^{+}08, LPZ08, LP08, LR05b, LS06c, LS10b, LHH11, MS12a, MWYF10, MBG12, MT09c, MRvM07, Mal08, MZ07b]. USA [MSZ08, MNZ09, MS10a, Mar11a, Mar11b, MRY^{+}11, MBP^{+}11, McG08, MW09, MMM10, MS96b, MNPW07, MZC06, MMO^{+}11a, MMO^{+}11b, Men05, Men07, MR08, MAFS08a, MAFS08b, MA11, MMG^{+}01, MCSY07, MGH09, MMK^{+}05, MNLM06, MS11a, MF06, MBA05, MAR05b, MCdS^{+}08, NNPW12, NP12a, NZS^{+}06, NSS^{+}07, Ngo09, NKN^{+}05, NQL06, NLS08, NS99, OGM05, OPS11, OR08, OSRS07, OPKR05, OZ09, OZ11, Pac05, PAV^{+}10, PM08, PHZ11, PDML11, PS09a, PZK07, PRK12, Per11a, Per11b, PCS^{+}07, Pfe06, PTM^{+}11, Pie10, Poi06, PS09b, PS07b, PISW05, PPM09, PM09, PT06b, Rab10, RMR05, RPR08, RMDS10, Raj09, RVY05, Rau11, Rei09, RBF06, Rob11, RL11b, RBsC^{+}06, Rog11, REPH06, RZ12, Sab10, SS07a, SNC12b, ŞS05a, SS11a, SYNC11, SS09a, SS11c, SW09a, STL11, SCHS07, SEG09, SF11]. USA [Sch01a, SS09b, SW09b, Sho05, Shu09, Shu11a, Shu11b, SPON08, SBG04, SS09c, SS11d, SSM08, SBM11, SH07, SR11, SF08c, Ste09b, Ste09d, Ste09c, Ste11e, Ste11f, Ste11d, Ste11c, SS09d, SBF08, SNSR06, SvASD05a, SvASD05b, SvASD05c, TT95, TD07, UDF06, VZ06, VH04, VS05, Wag08, WPR07, Wil99, Win09, WD10a, WD10b, XKGS12, XL11, YGE12, YDKM06, ZGK^{+}07, ZCRL08, ZSM^{+}09, dKIM^{+}07, vEHD05, vHH09, vdBGG10, vdMS11, AH00, AO02, AKZ01, ABS01, AYB95, Ada96, AH94, AS99, AH02, APSZ02, AL03a, AP04b, AP04a, ACK00, Ang97, Apo92, AJRS03, Auc98, Baa03, BCB99, BR03, Ban92, Ban93, Ban94, BG93a, BCT02b, Bel00, BMP90, BN04, Bla03a, BS94, Bon03, Bor94, Bri93, Bro92, Bro94, BCR03, BS01c, Bue04, Buh98, CP98a, COT98, CMT02]. USA [CF00a, CL01, CTT93, Cha02b, CPC^{+}99, Che03a, CR04a, CMZL04, CKO01, CW00, Cle92, CH98, CPvdH03, CS02b, Cop95, CM04b, Cou03, CG94, CT99, DW03a, DM01b, DSB04, DG02, DPB^{+}04a, DST01, Dei90, DDJ00, DS01a, Der89, Des94, Des02, DSM03, DH94, Die03, Dil94, DS03, DPP98, DKR02, DG97, Dwa00, EH95, EM02, Ehr87, EV01, Ell95, EKT94, EG97, ES00a, ET01b, FV01, FV03, FC98, FGH98, FK87, Fed01, FR95a, FR98, FR01a, FRS03, FSM03, Fer98, FNY93, Fos99, FT04b, Fre96, Gac02, Gae96, Gan96, GHR89, GMV03, GLP99, GHSW04, GJRT01, Goe98, GM99, GK02, GW98, GF97, Gup99, HI93, HPV00, HBS00, Heb95, HNM02, HM03, HS98a, HK01, HRT^{+}03, HF97, HJRS99, HS98b, HSB^{+}96, HH03, Hug91, HJ00, HS03]. USA [ID03, Ibr95, IL01, IWC04, Jaf99, JKRR04, JS01, Jay04, JSW03, JL00, Jen01, JP03, Joy03, JQ04, Jue04, KS03a, Kah87, Kal97, KKP02, Kap92, KT96, KP04, KS03c, Kil01, KD99, KK00, Kob96, KP99, KP01b, KMSS02, Kra98, KR02c, LLS00, Lam96, LSW02, LMS96, LG94, LS95, Lee01, LMW01, Lei95, Li98, LN03a, LG95, LPA00, Lom93, LHK02, Lov82, Luk97, LO88, LK00, Mai88, Mai90, MNT95, Mar01, Mar93, MNW01, MS00a, MTO99, MMR91, Mau89, MW03, McA00, MT02b, Mic90, Mon01, Mos01, MS02c, Mue01, Mur99, Nac01, Nao04, NW97, Nor04, Nor97, O’H98, Oka04, Pag98, Pal00a, Pal00b, PS97a, PS98a, Par02, PV02, PNB04, PN93, Pic01, Pin95, PC00, Por98, Pre02, PB92, RLL04, Ran93, RZ91, RZ94, RS97b, RO00a, Ras01]. USA [RA00, Ric02, Rol98, Rol99, RV02, Rus00, SW04, San02, SW03a, SW03b, SWAF95, Seh97, SdDK91, Sho84, Sil01, SRW98, SL99a, Sle92, SLLS03, SLF91, SC96, SWB03, Sti94, Str99, SV01, SC00, TDS94, Tom91, TG02, THR03, Van02, VKJ03, WZ03, WC98, WW02, WvdH03, Wie99, WGL96, WPD04, YD98, Yun02, ZZ04a, ZG03, Zip92, Zuc03, vEDA04]. USAB [Hol07, Hol08, HM09, HS11, LHH10]. Usability [Ayk07a, Ayk07b, Hol07, Hol08, HM09, HS11, Jac07c, LHH10, Mar11a, Mar11b, Ayk07a, Ayk07b]. Usable [DD07]. Usage [MNLM06, NZS^{+}06, NSS^{+}07, ZSSM03, ZSM^{+}09, MS00a]. USDM [BGG^{+}12]. Use [BAdVR08, CBF09, DZGS06, FD03, HP02a, HOC07, dVGR04, Ber98, FH01, FLS05, Her92a, KDS01, MMR93]. USEC [BDC12, DD07]. Useful [AGB^{+}11]. USENIX [Alo05, BC09a, CC07, GM10, IS08, Jac04, KK11e, vSH06, ES03]. User [AD98, AK12, Ayk07a, BHMS07, BM03b, BDN08, CHD07, CS03, CD01, CDND11, DJNR06, DGSN11, DKD05, DLM99, DKP00, DLO03, EPVW10, KGRD10, KKDV02, KKD04, KBC^{+}09, KCO04, LKD08, MMBND07, Mar11a, Mar11b, MTWD06, RWD09, SS04, SP07, TBD12, WJP07, Bee86, BDW97, DEM81, Hou90, JW74, JW75, MSRP98, MS00a, Mit98, ABM05, BGV01, BCR03, CMP07, CDFP11, DKC10, HMPZ09, KCMO11, MMDN12, PRdRW09, SS04].

User-Centered [SS04]. user-friendly [MSRP98]. Users [Ber02b, BBKW03, BH95, BHT97, BFH98, BDGK00, CLS07a, FT04b, Jac11d, JK06, Ste11f, SBH^{+}04, TKHS05, Ric02]. Using [Akm01, BST03, BCRB^{+}11, BSB^{+}99, CD07, DKLT05, KRRBV07, Mit06, Sub05, BST01, BM03b, Bla03a, Hai82, Hau86, LZ02, Rei93, Ruc96, Sch99, She85, San80b]. USM [LPH00]. USN [DsSL^{+}07, EYX^{+}05, ZSY^{+}06]. USSR [BBL87, MLM90, MT89]. UT [CMB^{+}08, FKM08, GQ11, GM11b, SS07a, SWAF95]. Utah [LLS00, LMT00, SBG04]. Utility [SW04]. Utilizing [Hof03a]. UTP [But10, DS06, Qin10]. Utrecht [AG01, ADP09, BEOT06, BG93b, BPD06, BDDS06, BCK10, BSH06, DJS06, DvEF06, EKO08, GM06a, HS95a, KBHSW06, LSJ06, Mad92, Mid01, MZ06a, NV01a, PMMR06, PLG06, PBS08, Pro96, SA06, TGH06, TtHVS06, WPM06, WBW94, ter97, vB97, vL94a, vL94b]. UUWSN [DsSL^{+}07]. Uzbek [EK81a].

V [ABG09b, BSvRW08, BKLM07, BCRT92, BT06, BLVV07, CGG^{+}07, GTWC09, GOZ05, LT05, OGM05, PAV^{+}10, BH77, AKL^{+}99, BLB^{+}03, BCT02c, DMP10e, EBSS98, FPRAW04, FK02, F^{+}12c, GGK^{+}06e, HKW12a, HZLL12e, JDK12, KZ01, LHK02, Mic90, MSR99, MGI^{+}11e, Ngu11c, PCMY11, PS06, PHPL06, RO09, Shi10, SAC^{+}06, dLDG^{+}08]. v.d.H [HS76]. VA [BKH^{+}09, CCJS07, CEER96, EK09, GP12, Li11b, PS09b, PPM09, Mar01, CMT02, CT99, Jaf99, PNB04, THR03]. VADER [MDH11a]. Vadhan [Gol11b]. Vadstena [LR94]. Vagueness [NvRSS11]. Val [CCR86]. Valbonne [BJN03]. Valbonne-Sophia [BJN03]. Valdivia [CHK06]. Valencia [AV08, AMM99, CGMP11, CvH07, DDHP05, DPMP^{+}11a, DPMP^{+}11b, EOE^{+}07, GB99, Han09, Hof03b, MMR07, Nie03, RV10, SSRA98, SG10]. Validation [CKLM09, MS10b, MS10c, MS12b, MSS12, BHH^{+}04a]. Vallée [BCP99]. Valletta [MC94]. Valley [SPÖ^{+}10]. Valparaiso [BYGP95, RMK07, GLMP07]. value [Chi79]. Vancouver [AW01, AFG^{+}11, BS11a, BM02b, BFKO06, CMQ96, DM01a, Deh99, DF11, HM05a, HV98, JS94, KV04, Ker01, MN98, MG07b, NBK04, PSS^{+}10a, SS01, STA^{+}11, TDK^{+}10, WGMP09]. Vancover [ST07]. VAPP [BCRT92, BV94, Bur90]. Varenna [ACP01]. Variability [DMB^{+}11]. Variable [TCYZ12, VZM^{+}10]. Variational [BtHRBB12, SMP07, TMLL09, PFCS05]. Variations [MO06]. Various [Jac09b]. Varna [CD00a, DD10, Dji89, DPT08, ED06b, INR07, RA12, Sco02]. Vary [MMP05]. Varzim [AMPT09, CK06b, CK06a, CK08, CK10a]. Varzin [CK10b]. Västerås [GHC^{+}06, HCR06]. VBC [HK96]. VCHCI [GT93]. VDAT [RCC12]. VDM [Bjo87, BHL90, BMJ88, DFG^{+}12, Bjo87, BHL90, BMJ88, PT91a, PT91b]. VDM-Europe [Bjo87, BMJ88]. VDMB [DKY06]. VECPAR [DDHP05, DPC^{+}07, DP97, PDH99, PAD^{+}08a, PDML11, PDH01, PDHS03]. Vector [BCRT92, BV94, Bur90, DP97, LV02, RR89, DP97, PDH99, PDH01]. Vegas [BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BFJM09, FILP06, GVO^{+}10, GSYZ10, GHP09, PS98a, PRK12, Ram01]. Vehicles [SFV^{+}11, VMB^{+}12]. Velden [EM03]. Veldhoven [ACvLdR03, RV98]. Venezuela [BV12]. Venice [ABBFM11, AGI^{+}06, BTT90, BBM^{+}08, BG09, Bil98, BB10, Bos00b, BPSW06a, BPSW06b, CNP^{+}08, CGGO03, CF99, Cor02, CCSW12, DLO03, KV08, KKL12, LB78, NR06, NSL09, PH97, PH11b, PB09, Pue07, RT09, SCM88, SL03, SGD09, TEB09, WKB04, Yah11]. Verbal [EVV^{+}11, EFZKM07, EBAH08]. Verbania [SPO11]. Verbania-Intra [SPO11]. Verification [AFKL04, AP04b, BJ06, BHKR11, BFF^{+}10, BM11a, BDG12a, BP09a, BZU07, BM09a, BC10c, BL02b, BH11, CJ11, CCK^{+}08, CH09, CBF01, DH07a, DB07, ELS12, FLUV02, GT03, GH06, GQ11, GZ06, GP08b, GM08b, HNRW06, HS02, JJF03, KS12b, KP15, LK99, Leu08, LR09, MS12a, MP10a, MS10b, MS10c, MS12b, MPS98, MN12a, MN12b, Mül02, NYHO07, NZZ11, ST07, TCJ10, UBW06, Wan04, Yor08, AH96, AHS96, Bar85, BP05, Bur97, CE95, Cas01, CM12, CB06, CK91, Cou93a, Dam97, DG01, Dil94, ES00a, ER05, God96, Gru97, HP99a, HPV00, Hoo91, HV98, HS03, Joh01, KS97, LS92, LB90, MSS12, MM01, MP93, MO97, PP01, PT05, Sif90, Wol95, vP93, BH10, BHS07, CP07, Cor02, Cou05, Der04, EN06, JS11a, JMO09, KR12]. Verification [LPZ08, SW10, SL03, Zuc03, Gre75, Pol81]. verified [Hun94, JMP12, LOR10, MW08, SW08]. Verify [CDE^{+}07]. Verifying [Hai82]. Verita [Nie94, GKOT00]. Vermont [DG97]. Verona [CSAH09, Eta05, Gia04, RtTMP10, SC09b]. Versailles [BKM01, BKZ05, GP05, GL74a, GL74b]. version [Koh06b]. versus [Len04, MB96]. Verwaltung [Sch74]. very [Hil78]. VHPC [ADB^{+}12a, ADB^{+}12b, BFT^{+}08, CAS^{+}09b, GVT^{+}11, LAF^{+}10]. VI [AG11, BSvRW09, BV94, BR07c, DFPV11, DGR06, FKP^{+}05, FRA05a, FNGW08, HKW^{+}12b, KOFJ09, MD04, MZ06a, PSD^{+}07, SA06, TT06b, Ang97, Bue04, DMP10f, F^{+}12d, GT09b, JL00, JvdAM^{+}12, KLR03, KOS02, Ngu12a, PCM11, PP06, SDR^{+}00, SEK^{+}11, SACM06, dLFG^{+}09]. Viareggio [BCG^{+}11]. ViC [NvRSS11, SCZ10]. Victoria [GN08, IS11b, IZ00, KL05, Yao95]. Video [BHL^{+}03, CL06a, CMOZM11, EKO^{+}04, Ho12a, Ho12b, KJR05, LSE02, MR07a, SBC^{+}94, SNSR06, WHL09, BCB97, BS01a, CKMS05, Her92b, KN03, LLC^{+}05, LG95, Ran93, ZVU99, CMOZM11]. Video-Based [KJR05, BCB97, BS01a, KN03]. Videos [ÜÇA10]. Vienna [AI06, AS07b, AKTT12, BBE01, BM03a, BBB^{+}07a, BP12a, BH06b, BCKR10, BBR^{+}10, BS02b, BKLR12, BNZ06, Bra93, CCF^{+}08, CCG08, CHR10, CD12, DF10, DR09, DHO08, DPSU11, DFT09, DBH01, DFS06, ED06a, EBT11, EFT01, EK12, FM04, FHKQ12, Fra00, FPRS06, GN90, GLM97, Gra05, GT93, GH93, HHMT12, HRdV11, HK94, HP06, HS97, HS08c, HHG11, Hut94, JP07b, JK05a, KR02a, KR04a, KJ09, KWC^{+}12, KLLP12, Kra03, KLN07, KSH05, KKH07, KS12d, LSTZ12a, LSTZ12b, LS06b, MKZK10a, MKZK10b, MJL02, MZ06b, PV07, PSD12, PGG^{+}06, PMDK00, PDG92, QSBC98, RCT05, RK06, SL12, Sch10b, Ses06, SS04, TR12, TQYX11, TBD12, Val06, ZBW94, BJ78]. vierte [Sie75]. VIETCRYPT [Ngu06]. Vietnam [BHH08, BOH^{+}12, GCSR07, HCL05, HZ08, HW05, JM02, Ngu06, NLŚ10a, NLŚ10b, NHJ12a, NHJ12b, RK09, TS10]. VIETRI [AMT03, MT02a, AMNT06, AHJ07a, AHJ07b, AHJ07c, EFZKM07, EHMM09, FSV09, GG98, MTV09]. VIEW [LLP^{+}07, STL11, CPRS01, Sch99]. Viewpoint [PR07]. Vigo [FHW12]. Vigoni [aCF06]. VII [AT07, BBvRL10, BSTL07, BS06c, CvRVSN12, HPP06, HSB^{+}10, IST07, KMK10, OROD07, PZ07a, PSM^{+}07, CdLG10, CL01, FLP^{+}12d, GT10, KOOL03, MAB^{+}02, Ngu12b, PCM^{+}12a, PBBY04, Por98, PIMN06, Shi12, Spa06]. VIII [ARSB09, APN08, AOSV08, LP08, NRSSV05, OSV11, SS08a, SK12, SSU12, YBL^{+}04, vdPS08, GT11a, KMSJ11, KOK04, MT02b, NRSSV05, Ngu12c, PCM^{+}12b, PS08b, Pri07].

Vilamoura

[MT07b, MT07c, MTH07a, MTH07b, MDH09a, MDH09b, MHD09, SdSMO12]. Vilanova [CPS05, MBVTY11]. Villa [aCF06]. Villach [HG00, LSK08]. village [Cho88]. Villard [War94a]. Villars [SF08b]. Villars-sur-Ollon [SF08b]. Villeneuve [GLM^{+}08]. Vilnius [CE01, Mit06, VMB^{+}12]. Virginia [Ada96, Ehr87, For88, KD99]. Virtual [ABB^{+}10a, BST03, CHD07, CD07, CHCM11, DKP00, DLO03, GYA^{+}06, KKDV02, KKD04, LKD08, MTT98, MTWD06, NNPW12, PMA^{+}07, PLI08, RABR03, RWD09, RKNV09, Shu11a, Shu11b, Sub05, TGM97, VKMT11, WK08, AD98, Aya95, BST01, Bod96, BKS96, BDW97, CD01, DAB01, DEM81, DKD05, DLM99, PGA^{+}05, BST01, BST03, CD07, Heu98, Heu00, Shu07, Shu09, Shu11a, Shu11b, Sub05, WKD08, YC11]. Virtualization [CGS07]. Visakhapatnam [PSDS11a, PSDS11b]. Visegrad [DT89, KN83, KSTP99, MTW05]. Vision [AMGJ12, AKR03, BCE09, BS06b, BTPPS06, BTPPS07, BTBP^{+}08, BTPPS09, BTBP^{+}10a, BTBP^{+}10b, BTKP^{+}11, BTPP^{+}12, BC10b, BKW09b, BTCW10a, BTCW10b, BTCW12, BKLS11, BtHRBB12, BLPW02, CV08, CN06, CBBS09, CDT11, DMP10a, DMP10b, DMP10c, DMP10d, DMP10e, DMP10f, DFP^{+}12, FLP^{+}12b, FLP^{+}12c, FLP^{+}12d, FTZ08a, FTZ08b, FTZ08c, FTZ08d, FSP09, FMC12a, FMC12b, FMC12c, GP07, GP09a, GP11, GVT08, HSNJ02d, HSNJ02a, HSNJ02b, HSNJ02c, HSL06, KP06, KKS11c, KKS11d, KKS11e, KKS11f, KHG01, KPS01, KH11a, KH11b, Kut12a, Kut12b, LBP06a, LBP06b, LBP06c, LBP06d, MRSV07, MLTC11, NNS06a, NNS06b, PM04a, PM04b, PM04c, PM04d, Pau01, RFZ03, SK11, SLH04, SMP07, SK08, SKK04, TMLL09, TZS00, TGM97, VHM07, YKKZ07a, YKKZ07b, YZCW07, ZiTM10a, ZiTM10b, ZiTM10c, ZJL06]. vision [Aya95, BTPPS05, BN98a, BN98b, BC96b, BC96c, CP98b, CP98c, CH99, Cro03, DDFM05, Ekl94a, Ekl94b, Fau90, FZJ01, F^{+}12a, F^{+}12b, F^{+}12c, F^{+}12d, FMdGC99, GL03, HP99b, Heb95, Ker01, KSW05, LBP00, Li96, MZF94, Nie99, OFK05, PTRH05, PFCS05, PH97, PZH96, RVY05, San92, SS01, SLH05, Ver00a, Ver00b, ZLJ05, ter97, TZS00, Sch99]. Vision/Computer [GP07, GP09a, GP11]. Visioning [LDC^{+}09]. Visions [WBHR08]. VISUAL [BL06c, BBP^{+}06a, BBP^{+}06b, BBP^{+}07a, BBP^{+}07b, BBP^{+}08a, BBP^{+}08b, BBP^{+}09b, BBP^{+}09a, BBP^{+}10a, BBP^{+}10b, BBP^{+}10c, BBP^{+}11a, BBP^{+}11b, BBP^{+}12a, BBP^{+}12b, CFH^{+}12, CAdB01, Fri06, HM12c, Lau00, LLP^{+}07, Nie09, QLXL07, QCDM06, SVS08, SBM08, ZRP^{+}09, ZRP^{+}11a, ZRP^{+}11b, BBKP06, CCL02, Cip95, CRYS09, Mac06, Ruc96, CCL02, Lau00, AGL06, BC02, BL06c, CAdB01, GMS03, HS99, Leu97, SVS08, SBM08, ZRP^{+}09, ZRP^{+}11a, ZRP^{+}11b]. VISUAL’98 [MSRP98]. Visualization [EDGP11, EH08, KEM07, KSFN08, Luo04, Luo06, Luo07, Luo08, Luo09, Luo10, Luo11, Luo12, TK05, DHM00, Die02, LG94, Luo05, MSRP98, OK90, WGL96, GCM08, GT87, GT90, HK96, TU94]. Vivid [Wag94]. VizSec [GCM08]. Vladimir [ELLP12]. Vladivostok [HM11c]. VLBV [AGL06, GMS03]. VLDB [CCG08, Col07, DW06, Jon01, JP04b, JP05b, JP06, JP07b, JP08, JP09b, JP10, JP11b, JP12a, Pie05, Bre03, Jon00]. VLDB-99 [Jon00]. VLSI [RCC12, Kun84, Mak86, Rei88, Van85]. VLSM [PFCS05]. VMCAI [BH10, CP07, Cou05, EN06, JS11a, JMO09, KR12, LPZ08, Cor02, SL03, Zuc03]. VMR [Shu09]. VOCUS [Fri06]. Volas [BH05]. Volterra [KL92]. Volume [BFMR96a, BFMR96b, Buc85, BN98a, BN98b, Cav85, CP98b, CP98c, DO89a, DO89b, EEK^{+}87, EKLM87, AM91b, AM91a].

Voronoi [GTA10, GTM11, Kle89]. Voss [TMLL09]. VOTE [AV07, RS09d]. VOTE-ID [AV07, RS09d]. VoteID [KL12]. Voting [AV07, CJR^{+}10, KL12, RS09d]. VSMM [WKD08, ZPT^{+}06]. VSTTE [JMP12, LOR10, MW08, SW08]. VT [Bue04]. Vugh [GKST11]. Vulnerability [BL06d, FB09, HS07, HB11, KJ10, Zam08, JK05a]. VW [Heu98, Heu00].

W2GIS [BRL08, CT06, CFM09, DPT12, KBC05, LV05, TFK11, WT07]. WA [BJ06, BG06, CJ11, DPB^{+}04b, ET06, FS08, FS06, GA09, HNRW06, JP11b, MW09, MNLM06, NP12a, Pfe06, VZ06, Bor94, DSB04, DPB^{+}04a, DSM03, FRS03, LMS96, MT02b]. WAA [TWYhL01]. WABI [BM06, CL08b, GH07a, MS10e, MC05, PS11b, RT12, SW09a, BP03, GM01, GG02b, JK04]. WAC [Smi05, SS06d]. WACG [LM96]. WADGMM [KMS12]. WADS [DSS91, Deh93, DLOS05, DSZ07, DGST09, DIS11, Akl95, DSS89, Deh97, DST01, DSS03]. WADS’99 [Deh99]. WADT [CM09c, FMO05, FS07c, MK12, CR01a, Par98, WPH03]. WADT’98 [Fia99]. WADT’99 [BCM00]. WAE [BFMS01, NW01]. WAE’99 [VZ99]. WAGA [DJ90]. WAHC [CSN14]. Waidhofen [MPPU11]. Waidhofen/Ybbs [MPPU11]. WAIFI [CS07, HH10, ÖRH12, vzGIK08]. WAIM [BGG^{+}12, CWC^{+}07, CLZ^{+}09, CTYG10, DLW^{+}07, FWY05, GLW^{+}12, LFP^{+}09, SPÖ^{+}10, WLO^{+}11, WJL^{+}12, YKL06, DTW03, LWF04, LZ00, MSW02, WYL01]. Wakamatsu [Bha05, Bha07, KSB10, KMSB11, OS03a, WVDS05, YGGJ04, GY03]. WALCOM [DU09, KK11b, iNR08, RF10, RiN12, DU09, KK11b, iNR08, RF10, RiN12]. Wales [Emb98]. Walther [SW10]. Walworth [DS06]. WAOA [BS09, BJ10, EP06, EK07, JSO11, KS08a, PSO05, SOP12, JSO04].

WAPCV [PTRH05, PR07, PT09]. ware [BLS02]. Warehouses [Jür02]. Warehousing [CD11, CD12, Kam99, KMT00, KMW03, KMW04, PMT09, PMT10, SEN07, SEN08, TT06a, KWA01, KWA02, MT99b, TT05]. Warming [AKTT12, KT11]. Warrenton [Ehr87]. Warsaw [BDIR07a, BDIR07b, Bih03, Bli75, BMN05, BKW09b, BTCW10a, BTCW10b, BTCW12, BKL^{+}12c, BKB03, Deg03, DR02b, DKOZ05b, DKOZ05a, GH03, Gor03, Hed03, KKdM^{+}07a, KKdM^{+}07b, KPRS07, KRSR11, MW97, MS11b, Pez03, PS98b, RS99b, RTZ08, Ska01, SKR^{+}10]. Warwick [CMPW12a, CMPW12b, GM89, Jos88, Pat90, PC99]. WASA [CLZ06, CEQ^{+}11, LHDD08, LBDW09, PKM^{+}10, WZJX12]. Washington [AKZ01, BS94, BS01c, Can12, DHRT10, Fei03, FT04b, GP09b, JS01, Luk97, OR08, PS07b, SW03a, SF08c]. Waterloo [BGS11, BAP^{+}05, CM09a, DLOS05, FHH11, HH04, HJ12, ST01]. Watermarking [HSKB09, KCR04, KKH09, KSB11, SJ06, SKK08, SKPG12, BCKK05, CKL05, PK03]. Wavelet [TWYhL01]. WAW [A^{+}08, ADL09, BC07, BJ12a, FHP11, KS10b, Leo04]. way [BMJ88, Zui99]. WBIR [DCFR12, FDL10, PLG06, GMV03]. WBL [LWM^{+}08]. WCC [Ytr06]. WCCI [LABM^{+}12, ZYW08]. WCII [KIRK05]. WCMT [CLZ^{+}09, SPÖ^{+}10]. WCNS [CGMP11]. WCW [CvSW04]. WDAG [BM96, HR95, MT97, Sch93b, SZ92, TV94, TSK92]. WDPP [CLLD09]. WDV [VHM07]. WEA [AS06, Dem07, McG08, Nik05, Jen03, RM04a]. Web [APR^{+}06, ACN^{+}07, ABG^{+}06, AST05, ABMS06, AB04, AAB^{+}07, AGS^{+}11b, AGS^{+}11a, ATC^{+}09, AAH^{+}10b, AAH^{+}10a, AWA^{+}11a, AWA^{+}11b, ABW10, AMM99, BBF^{+}06, BBM^{+}08, BGP08, BMP00, BMP01, BTQ02, BTQ03, BBP04, BPW06, BHHK08, BCG^{+}07, BCBG05, BHM^{+}04, BHMS07, BPZ09, BKH^{+}09, BG12b, BJ12a, BKZ05, Bre03, BSS00, BKN07, BHM03, BM09b, BHM^{+}02, BDFS04, BFOY04, BTF05, CS05a, CPSA10, CB12, CWC^{+}07, Cha03a, Che99a, CLZ^{+}09, CCG08, CDA^{+}06, DF10, DN04, DHRT10, DB09, DCMS09, DA08, DTW03, DLW^{+}07, DFW^{+}11, EM05, EK12, FS05, FWY05, FMC04, FSM03, FCLL02, FKM07, FHP11, GCFA12, GMBM05, GPB02, GPE05, GPYD09, GSV05, GHI^{+}06, HK12, HH02, IHX^{+}08, JLST06, Kro00, KS10b, LLCL05, Leo04, LLLL08, LPC^{+}11, LWF04, LZS^{+}08, LFP^{+}09, LMT00, LSL04, LLL06, LSW^{+}10, LCY^{+}11, LDC^{+}09, MPvH04]. Web [MSS03a, MSW02, MSG06, MA05c, MNLM06, MSS04, NZS^{+}06, NSS^{+}07, NKPH08, NKN^{+}05, NKC09, OS04, OPS11, PCK^{+}12, PSPH^{+}10a, PSPH^{+}10b, Paz03, Per06, PKSS02, PdA^{+}11, PLK^{+}12, PBW05, PW07, PW08, SW03b, SVS08, SLL^{+}06, SWJX12, SSD^{+}08, SCP^{+}12, SLKL09, SV01, SKN05, TFE^{+}09, Ull08, WAS06, YLLZ04, YKL06, ZSSM03, ZTY^{+}05, ZYBX08, ZSM^{+}09, ZNCF03, ZZO03, ZLS^{+}06, dCdF^{+}08, dBBC02, A^{+}08, ADL09, BC07, BMP01, APR^{+}06, ADP11, BMS^{+}08a, BGG^{+}12, BFH07, Bar03, BTQ02, BCG^{+}07, BCKR10, BHM^{+}04, BRL08, BHL11, BTT12, BNZ06, BB11, BW09, BHM^{+}02, BkHJ^{+}04, BFOY04, CL08a, CP12, CT06, CFM09, Cha03a, CTYG10, CTS10, CvSW04, CBS^{+}11, Chu03, Cue03, DGJ^{+}05, DPT12, DCMS09, DFW^{+}11, DH08, FWZH06, GGD09, GLW^{+}12, GLC^{+}11a, GLC^{+}11b, GCC^{+}02a, HZK08, HL10, JZ03, KFW04, KS12d, KBC05]. Web [LS10a, LV05, LLWL09, LG05, LZ00, MPdSM07, MS00a, MD01, NKN^{+}05, NM97, PS09b, RG11, SPÖ^{+}10, SWJX12, SILS05, SD06, TFK11, Vis04, VLY09, WYL01, WGLL10, WLO^{+}11, WJL^{+}12, W^{+}12, WZH^{+}12, WT07, WHG07, ZJ04b, ZTY^{+}05, Zho01, ZLY^{+}07a, ZZO03, ZSP^{+}04]. Web- [Cha03a]. Web-Age [DTW03, DLW^{+}07, FWY05, LWF04, MSW02, YKL06, BGG^{+}12, CTYG10, GLW^{+}12, SPÖ^{+}10, WLO^{+}11, WJL^{+}12, LZ00, WYL01]. Web-Based [BSS00, FCLL02, LPC^{+}11, LSW^{+}10, LCY^{+}11, NKPH08, PLK^{+}12, SVS08, Ull08, WAS06, ZNCF03, DN04, LLCL05, LSL04, dBBC02]. Web-Graph [KS10b, Leo04, A^{+}08, ADL09, BC07]. Web-services [Cha03a]. WebDB [SV01]. WebDB’98 [AMM99]. WebETrends [CWC^{+}07]. WebKDD [MNLM06, NZS^{+}06, NSS^{+}07, ZSM^{+}09, KMSS02, ZSSM03]. WEBKDD’99 [MS00a]. WebMine [BHMS07]. WECSR [BDC12, DDS12, SCD^{+}10]. Wedekind [HL05]. WEED [VMvN12]. Weihai [Che10, SLZH07]. Weimar [AL12]. WELCOM [FMW01]. Well [IdKM06]. Well-Being [IdKM06]. Wellington [Kie12, NHJ04a, NHJ04b, NHJ04c, WCK11]. Wellness [MKB^{+}09]. Wendisch [BJ86a]. Wendisch-Rietz [BJ86a]. WES [BHM^{+}02, BFOY04]. WESOA [MRY^{+}11, PJC^{+}12]. West [DR06, EV01, ELM05, EdS08, Fra01, Fra99, Hir97, Hir98, Jue04, Nic75, PH04, Syv02, TTF07, Wri03]. Western [AL12, LSW08]. Westfalen [BHR84]. WEWoRC [AL12, LSW08]. Weybridge [Gru98]. WFLP [Esc10, Kuc11, Mar11c]. WG [AK10, AZ11, APP12, Atl08, BBdB12, BA07, BdB08, BJ07, BP05, BKM07, BGR08, BEFP08, BD11, CKD11, CK12a, CBCGA12, DL06a, DV07a, DFPS06, EDGP11, EM06, EK10, FR11, Fom06, FJ10, GR12, GLIC10, GSLM12, GW06a, GH12b, GS08, GV09, HZ10, HNW04, HHG11, IR07, JW05, JSWhT11, KD05, KK11d, Kra05, KA05, LLPH09, LN09, Li11b, MBH^{+}09, MT08c, MW08, MPPU11, NPPDG06, NBM09, ONPR07, OSC^{+}08, PVW10, PH10, PSM10, Pro11, QBY^{+}12, SNČ^{+}12a, STP^{+}10, SMBQ07, SJW^{+}12, SO09, SZ05, SHYEF08, SHUH08, SV12, TMG10, TMdB11, TMS12, Thi10, TQYX11, UDF06, Wan05a, WCJS10, WZ11, BS92, Bod03, BW00, BL01, CE95, CR01a, DO02, GW96, GT03, GS88, KHW03, Kuc̆02, KW96, MM01, May93, MST95, MPS93, Nag90, Nag95, NDFMS97, NNS03]. WG [Nol81, PV02, TS87, dFEN04, vL89, vL94b]. WG10.2 [MP93]. WG10.5 [LK99]. WG6.1 [PVTG07, SDH03]. WG’90 [Moh91]. WG’97 [Moh97]. WG’98 [HS98c]. WG’99 [WNE99]. WGIM [WJL^{+}12]. WGLBWS [LCY^{+}11]. Where [FMPT10]. While [GKL09]. Whistler [MR07b]. Whittlebury [MSS04]. Whole [Lam00]. WI [CvRVSN12, JSV11, CR04a, Zho01]. WI-IAT [CvRVSN12]. WIA [WY98]. WIA’96 [RWY97]. WIA’98 [CMZ99]. WIA’99 [BJ01a]. WICI [ZLY^{+}07a]. Wide [AMM99, BBB^{+}85, Che99a, LMT00, NM97, SV01]. Wien [Kui92, KRS02]. Wiesbaden [May93]. Wiesbaden-Naurod [May93]. Wigderson [Gol11b]. Wildbad [BEM91, Hac75]. WILF [BPT06, DPP09, FPP11, GMP06, MMP07]. Wilhelm [RSB07]. Wilheminenburg [STT04]. Willem [DHS10, MvOvRdV05]. Willem-Paul [DHS10]. Williamsburg [CEER96, KD99, Mar01, PPM09]. WImBI [ZLY^{+}07a]. Winchester [CKC11]. Windsor [Ber08, BF10, CAD^{+}02, MTNL12]. WINE [CEK11, DG07b, Leo09, PZ08, Sab10, SMK06, DY06b]. Winning [Lib07, Gur04]. Winnipeg [DS11c]. Winter [Bjo80]. WIOSCA [VMvN12]. Wired [BMK08, BMK09, BDK10a, BKH11, Cho02a, KHI06, KHS07, Tha84, BMM07, BCKT05, BCF06, HHLS08, KMMT12, LLMT04, MBVTY11, OKBMB10, PB09, vdBHOS09]. Wired-Wireless [PB09]. Wired/Wireless [BMK08, BMK09, BDK10a, BKH11, KHI06, KHS07, BCF06, HHLS08, KMMT12, LLMT04, MBVTY11, OKBMB10, vdBHOS09, BMM07, BCKT05]. Wireless [ASE^{+}07, AZ11, BMK08, BMK09, BDK10a, BKH11, BRL08, BPS06, BAP^{+}05, CT06, CFM09, Cho02b, CK06c, CGGO03, CSRS08, COB06, DPLW08, DPT12, Dol09, ENO12, Fek08, FLR11, GCC^{+}02b, KHI06, KHS07, KO07a, KR06, KCK08, LPR12, MVCM10, MKR^{+}04, NdG04, NBK04, NW10, NR04, NR06, PBK03, RGLA09, Sch10a, SWB06, TFK11, WT07, DB02, KRMM^{+}02, KBC05, LV05, Omi00, SC05, BCL03, BCF06, CA08, aCF06, CLZ06, CEQ^{+}11, GVCA07, HHLS08, KWW04, KS05b, KMMT12, LV07a, LLMT04, LHDD08, LBDW09, MW11, MBVTY11, OKBMB10, PKM^{+}10, PB09, PH12, RS09b, RKM06, SKB10, Ver08, WZJX12, vdBHOS09, BMM07, BCKT05]. WIRN [AMNT06, AMT03, MT02a]. Wirtschaft [Sch74]. WISA [CPGL05, CSY09, CY11b, JY12, KYL07, LYY07, LY12, LY05, SKY06, YY09, CY03]. Wisconsin [Pag98]. WISE [APR^{+}06, BMS^{+}08a, BCG^{+}07, BHL11, BkHJ^{+}04, CTS10, CBS^{+}11, DGJ^{+}05, FWZH06, HZK08, NKN^{+}05, VLY09, WHG07, ZSP^{+}04]. Wisepersons [Fur95, FU96]. WISH [THX^{+}07]. WISI [CWY^{+}06]. WISM [CVLL12, DMB^{+}11, GLC^{+}11a, GLC^{+}11b, LLWL09, SPC^{+}08, TDK^{+}10, WGLL10, W^{+}12]. WiSME [Bru06b]. WISS [CBS^{+}11]. WISTP [AZ11, APP12, MBH^{+}09, OSC^{+}08, STP^{+}10, SMBQ07]. Within [LSR03]. WITS [AL11b, DV09]. Witten [Amb01]. Wittenberg [ZT04]. WLC [SCD^{+}10]. WLP [SHGB05, SHW09]. WM [ADB^{+}05]. WMC [CFP^{+}09, EKP^{+}07, FPRS06, HPRS06, MPPJ^{+}05, PPJRN^{+}10, MVMP^{+}04, Pau03]. WMC-CdeA [Pau03]. WMS [MT^{+}03]. WNAA [VWY97]. Wokingham [BF06]. Wolfville [Tom92]. WoLLIC [BdQ11, DdQ10, HdQ08, LdQ07, OdQ12, OKdQ09]. Wollongong [CS12c, KjHGG11, MSS08, PSN95, PS96b, PSNS99, POS00, SNS03, SLM07, SMS12].

WOMP [M^{+}06]. WOMPAT

[Cha05b, EV01, Vos03]. WONS [BCL03]. WOPPLOT [BE87, BEM91]. Worcester [KP01b]. Worchester [KP99]. Word [AP93, Sch92]. Words [LSR03, NP85]. Work [CSPC09, CRJ11a, CRJ11b, Dai07, FSdlPT11a, FSdlPT11b, Hol08, Kam99, Kar09, LHH10, MMMD97, MSA99a, MSA99b, MP01a, MP01b, MA03b, MA03c, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, ORB^{+}09, Rob11, SPCG07, SCL^{+}06, SLL^{+}07, SYY^{+}08, WPD04, Bjo87, PSD^{+}07, PSM^{+}07, SLB05, MP94]. Work-Conference [CSPC09, CRJ11a, CRJ11b, FSdlPT11a, MMMD97, MP01a, MP01b, MÁ05a, MÁ05b, MÁ07a, MÁ07b, MFÁ^{+}09a, MFÁ^{+}09b, ORB^{+}09, SPCG07, FSdlPT11b]. Workflow [Rei03, Sch96, Sch98a]. Workflows [PF07b]. Workgroup [Hol07, Hol08, HM09, HS11, LHH10]. Working [Atl08, BA07, BPR05, BCL03, BF11, Bis02, DL06a, EW04, FJ10, GT03, GH09, GV09, GHP^{+}08, HDLM09, JW05, LM98a, LdRV94, LK99, LPH00, MM01, MNW01, MP93, MCSY07, MT99c, PR08, PVW10, RKS89, RD12a, SPH07, STTP02, Tah09, WSS04, WP10, Wim03, Wim04, Yas00, BP05, CE95, Chi79, DE05, PGA^{+}05]. Workshop [AK10, AP93, AKK04, AGP08, ABB^{+}11, AS99, AF97, ABF^{+}07, APS08, Akl95, AG98, Alb99, AJM87, ADP09, ABBC10, APO96, ABMS06, ACW07, AW90, AAB^{+}06, ACJ09, ANR99, AM07, AFM01, AC09, ANF08, ABR09, AR10, AP04a, Ama08b, ACK00, Amb01, Ame91b, And94, And96, AH98, ADMH04, ADM^{+}06, ADM^{+}08, AB04, AKL^{+}99, AT07, APN08, AMT03, AMNT06, AHJ07a, AHJ07b, AHJ07c, AFR^{+}04, AZ11, AF06, AJ94, Arm02, AL11b, AL12, AI98, AJRS03, AOSV08, APV09, AM02, AKR03, APP12, ABRZ06, AJ01a, AMM99, AL06, Auc98, AKS09, Avg12, AGCP^{+}11, ABRS06, AS07c, ABFR07, BM03a, BM96, BBMB^{+}08, BBD^{+}10, BCBF08, Bak95, BCB99, BEOT06, BE06, BSvRW08, BSvRW09, BBvRL10, BR03, BdBC01, BS09, BJ10, BFGM05, Ban92, Ban93, Ban94, BDZ04, BP04a, BL12, BP12a]. Workshop [BN94, BK05a, BBA12, BHJKPG06, BGH06a, BDE12, BG09, BBAB^{+}07, BRS^{+}07, BHP95, BEM91, BWAM09, BBDS12, BS06b, BC92, BdQ11, BBB07b, BS03a, BKR09, BDG^{+}12b, BP09a, BP03, BC96a, BCD04, BDW08, BHMS07, BS92, BMS04a, BR89, BCM00, BCV06, BDSW09, Ber03, BR97, Bet83, BCI85, BB89, BG11b, BM99b, Bha00, Bha02, Bha05, Bha07, BB03, BC93, BN91a, Bih97, BY07, BD98, BMS95, BGS11, BBH01, BS10, BBRS05, BBC^{+}08, BBD^{+}09, BBR^{+}10, BJN03, BPT06, BK02, Bod03, BL06b, BRU06a, BJS99, BU01, BHU03, BMN05, BS94, BJ01a, Bv97, BJ12a, BB06, BFMS08, BFJM09, BF89, BDDS05, BDDS06, BDDS07, BBR88, BBR89, BBR90, Bor92, BGR03, Bor94, BAL88, BL02a, BE07, BGJ11, Bos00b, BR07a, BM94, BDN08, BRM06, BP09b, BEM06, Bra02]. Workshop [Bra09a, Bra99, BW00, BL01, BKM07, BBT10, BKZ05, BNZ06, BB11, BHV11, BHH^{+}06, BFvdM10, BGHS05, BKNS06, BMS01, BAdVR08, BHLvdP07, BBH08, BGR08, Bri95, Bri97, BGY96, BFMS01, BEFP08, BR98, BKM06, BKM08b, BSH06, BHJY07, BW09, BK04b, BHM03, BM06, BCF^{+}02, BD08, BS01c, BBB06, BLPW02, Bun96, BKB03, BY06, BHM^{+}02, BFOY04, BCDN07, BMP96, BGH06b, BGW06, BB00b, CPS05, CJ11, CG12, CLMP02, CCS04, CPR^{+}10, CKP^{+}12, CCJS07, CKD11, CL11, CK12a, CV04a, Can12, CGL^{+}09, CBG^{+}10, CBS^{+}12, CV08, CMSM^{+}10, CP12, CS03, CS05a, CZ02, CS07, CBF09, CF00a, CGS01, CW94a, CM95, CL01, Ceb08, CA08, CR01a, aCF06, CZ05, CY03, CMZ99, Cha02a, Cha05b, CZG^{+}08, CGKM11, CMMR12, CJRR07, CPC^{+}99, CWZ^{+}11a, CWYC12, CP88, CLL^{+}09]. Workshop [CGL^{+}11, CJRT05, CR04a, CWY^{+}06, CYCL09, CF09a, CChL^{+}10, CEK11, CD00b, CL04, CvSW04, CM03b, CM09a, CHZ08, CW94b, CKO01, CS10, CBN99, CCLR98, Chr99, Chr00a, Chr01, CCMR02, CCMR03, CCMR05, CCMR06, CCMR07, CCMR09, CCMR10, CMMR11, CCMS11, CMSA12, CCG08, CSY09, CY11b, CNY95, CW01, CM09b, CFVCF^{+}12, CHD98, CER79, CG09, CGS07, CS00b, CH98, CPvdH03, CF09b, CBT07, Coh94, CMS11, CDN12, CR01b, CLS07a, CM00, CDW86, Con97, CKOB11, CMCS11, CDNS00, CAP08, CAdB01, CS97, CFP^{+}09, CM09c, Cor02, CKTW12, Cou93b, Cov99, Cra08, CL08b, CDM04, CRW^{+}12, CRS00, CH01, CLBP11, CEER96, DLFP02, DR02a, Dal92, DKY06, Dam97, Dam91, DM06, DG06a, DD97, DB02, DD03, DSDS04, DU09, DDE04, DSLT08, DSRW08, DSLT10, DSHL11]. Workshop [DH01, DS09a, DMO98, DM01a, DM01b, Dav91, DCFR12, DV10, DdQ10, DG01, DAB01, De 95, DKK08, DW12, DMW00, DGM09, DV09, DG10, DEG11, DSS89, Deh97, Deh99, DST01, DSS03, DSZ07, DFP^{+}12, DP11, Dem07, DB98b, DG07b, DBB12a, DL95a, Des02, DJS06, DJNR06, DLN10, DGSN11, DKN^{+}11, DSV01, DPP09, DP10a, DGWB09, DOS99, DJRZ06, DD07, Die03, Dig04, DGR06, DGDM^{+}11, DMRS06, DMRS07, DZGS06, DS03, Din03, DLdS04, DMB^{+}07, DKN12, DJNR09, Dit88, DPP97, DdCF98, DPP98, DL04b, DFN10, DLGJ10, DB07, DSL08, DHO08, Dol09, DO09, DO11, DFT04, DPSU11, DP12, DW94, DMW96, DMW06, DDS02, DGB^{+}06, DHMM96, DFD04, DW06, DTF98, DKR02, DSS^{+}07, DH08, DA01, DKN02, Dun09, DH07b, Dur12, DFGN12, Dwa00, Dwy01, DNS95, DF99]. Workshop [DS07, EDGP11, EKS10, EL07, EKF12, EM08, EKO08, EGO09, ENR83, Ehr87, EKR91, Ehr91, EO94, EEKR00, EV01, EdS08, EKR10, EKP^{+}07, ERK07, ET01b, EPVW10, EP06, Erl06, EK07, Esc10, EV07, Ési06, EFZKM07, Eug09, FRS08, FBKS03, FBSMS08, FL00, FV01, FV03, FV04, FV05, FPFR05, FPP11, FS07a, FD03, Fed01, FR97, FR98, FR00, FR01a, FRS02, FRS03, Fek08, FS99a, FKB02, FR95b, Fer96, FMZ05, FFL^{+}12, Fia99, FS07c, FKM09, FdCLN^{+}11, FMW01, FZJ01, FLP12a, FDL10, FO95, Fis08, FST09, FvdTDG12, Fle99, FR99a, Fog94, Fog95, Fog96, Fom06, FLUV02, FP08, FS07d, FS08, FS09b, FS10, FR99b, FRSH05, FKM08, Fre98, FPRS06, Fri01, FHP11, Fri93, Fuc98, Fuc10, FLMK06, FLT07, FKLT08, FK09, FCDB07, Fur95, FU96]. Workshop [GVO^{+}10, Gab13, Gae96, GKL09, Gal97a, GMV06, GMV07, GWY99, GPCL10, GANACBR10, GANACL11, GANACBdCdV12, GAL12, GVCA07, GMS03, GB99, GM01, GHR89, GMV03, GB97, GFV09, GKS^{+}07, GMV09a, GKMV10, GKMV11, GKMV12, GMP06, GW03, GKT09, GKT10, GKST11, GKS12, GG02a, GH07a, GCK06, GKL^{+}11, GK90, GH05, GSYZ10, GH06, GHI^{+}12, GPM98, GMO03, GHSW04, GLM^{+}08, GOW03, GOZ05, GGS11, GM06a, God05a, GP98, GJRR08, GJRT01, GJRR11, Gol96, GGHS07, GSLM12, GCM08, GM99, GSP01, GPS03, GLS05, GKS05, GZSC07, GS88, GN90, GGS99, GR06b, GM06b, GN05, GK09, GM04b, GP08b, GW06b, GNTF02, GM11b, GN08, GF12b, Gru98, GO06, GH93, Gud08, GAR03, GAR04, GRR05, GS06, GB07, GG02b, GJTS06, Gün88a, Gup99, GJRS12, GKOT00, HP02a, HOC07, Hag02, HK04]. Workshop [HO03, HKR07, HENR03, HG02, Hal00, HPTV05, Ham89, Ham91, HDC99, HY12, HP99b, HV03, HWW^{+}10, HAB02a, Har01, HS94, HG96, HK98a, HH10, HKM95, HPV00, HMP08, HOD96, HKS06, HHP09, Heb95, Hee94, HR95, Hel94, HM03, HS98a, HK01, HS02, HM02, Her92b, Her91, HT06a, HA00, HRT^{+}03, HRTR05, HRR^{+}06, HPS09, HR08, HCF06, HSKB09, HJRS99, HAB02b, HdQ08, HhKRT06, HKP11, HI10, HPRS06, Hor00, HMG03, HT06b, HS98c, HNW04, HSB^{+}96, HSL06, HKM^{+}07, HS08c, HHG11, HM98, Hut94, Hut96, HK07, HSTU99, HdMKS12, IMW04, IMK06, IS11a, IS11b, IZ98, IZ99, IZ00, IST07, IS98, Ish99b, IGN05, IFV07, IH90, IY95, IHQ96, IS12c, IN11, JRSN09, JMBS07, JR98, JK00, JLV02, JSO04, JKRR04, JSO11, JP07a, Jan89, Jan92, JLST06]. Workshop [JSW03, JL00, Jen03, JCC^{+}12, JFT11, Joh03, Joh01, JK04, Jon00, Jon01, JP04b, JP05b, JP06, JP07b, JP08, JP09b, JP10, JP11b, JP12a, JS02, JJF03, JDMB08, JDMB10, Jou11, JS94, JQ04, JP12b, JY12, JRR03, Kåg98, KEDW07, KS03a, KS09a, KS08a, KKP02, KCR04, KM08a, KK97, KR10a, KO91, KWW04, KL88, KL96, KS03b, KMD02, KT96, KMC^{+}01, Kat99, KK11b, KGL04, KS09b, KM08b, KJ91, KMF06, KSB10, KMSB11, Kim01a, KYL07, KKH09, KSB11, KM04, KDS01, KR00a, KR00b, KR01, KHG01, KPS01, KŽ04, Klu97, KW98, KSW99, KK00, KZ01, KOS02, KOOL03, KOK04, KRP06, KHPS07, KPP08, Knu99a, KP99, KP01b, KNP01, KvG98, KMSS02, KK09a, KK11d, KBHSW06, KHSM^{+}08, KS05a, KRMM^{+}02, KC00, KW10, KS07a, KS07b, KBČ^{+}12, KMS12, KGS08, KHK^{+}08, KVMH09]. Workshop [KR10b, Koz82, KH11c, Kra03, Kro00, Kru08, Kuc̆02, KP10c, Kuc11, KF12, KH12, KS10b, KW96, KCO04, KMR12, KCK08, KL02, LSJ06, LH11, LJS08, Lac10, LV12, LRS03, Lag05, LS10a, LSW00, LSW01, LSW02, LSW03, LO11, LN04a, Lau01, LS04a, wLN99, LBP00, Lee01, LV02, LB03, LSgL^{+}06, LYY07, LN09, LCY10, LMMN10, LY12, Leh11, LOST04, LT05, LOTY05, LTÅ^{+}11, LdQ07, Leo04, Leo09, LO98, LNE06, Leu08, LS05, LM08a, LWM^{+}08, Leu03, LPZ12, LLP^{+}07, LSHB07, Li98, LSDN04a, LSDN04b, LST^{+}05, LOS05, LBO05, LYT06, LEP^{+}10, LM96, LMP04, LMCH12, LYCV08, LG95, LL06, LLO08, LSIT11, LK04, dVLKR^{+}08, LK07, LK09b, Lom97, LN04b, LOH05, Lop06, LSF07, LH08, LHK02, LSU07, LRS98, LP08, LGS09, LSW08, LR05b, LCG97, LS09, LK00, LIH99, MWD06, MPP94, MPK04, MDY^{+}10, MDHB11]. Workshop [MKPV05, MMD05, MTT98, MT09b, Mag03, MP06, MBG12, Mak86, Mal97a, MP03a, MJ04, MZ99, MBT^{+}03, MST12, MS10a, MMBND07, MT04a, MS96a, MT02a, Mar11c, MBH^{+}09, MVCM10, MB01, Mar03, MVMP^{+}04, MOFM10, MS05, MP10b, MMP07, MMOS99, Mat02, MZ04, MF08, MPR07, MZCC09, MT97, May93, MST95, MRPM10, MRP11, MPR12, McG08, MB06, MMM10, MFG12, MNPW07, MC94, MH04, MLTC11, MT02b, MB09, MMG^{+}01, MMMO96, MMU96, MPPU11, MCP93, MS95, MV12a, MKR07, MMK08, MA05c, MNLM06, MP12, MOT12, MK01, Moh91, Moh97, MTW05, MvL10, ML95, MF06, MG07b, MDP04, MDBF01, MK03b, MO05b, Mos01, MK12, MS10e, MS02c, MB98, Mue01, Mug97, MWJ97, MSR99, MZ06a, MdSC09, MGSM12, MNW98, NP02, NL01, Nag90, Nag95, NDFMS97, NSM00, NW01]. Workshop [iNR08, NcC99, Nar01b, NZS^{+}06, NSS^{+}07, NZ06, NV08, NW97, NT98, Nie94, NR04, Nik05, NR06, Nol81, NSFB10, NS99, NvRSS11, NBM09, NSL09, NU92, ND04, NH03, Nyb08, O’H98, OROD07, O’S03b, OB07, ONPR07, OFR09, OLL10, OS04, OS09, Ölv10, OTZ01, OPT01, OPP04, OSV11, Omi00, Ong05, OdQ12, OSC^{+}08, OKdQ09, Oqu03, OWM04, OK96, OR08, ÖRH12, PSP^{+}02, PZ07a, PM08, PWZ98, PV07, PKKG06, PP01, Pal01, PR07, PT09, PPCS12, PS97a, PZ08, POP09, Par02, Par98, PMÓ06, PMMR06, Păs09, PMH09, PS12a, PH10, Pau03, PPJRN^{+}10, PW92a, PW92b, PZ96, PW09, PH97, PH11b, PA03a, PvES12, PS02b, PH02, PD04, PpWR96, PP99, Per01, PM96, PSD12, PBGK02, PBGK03, PCG^{+}05b, PGG^{+}06, PCG^{+}07, PJM^{+}08, PDF^{+}09, PDK^{+}10, PCG^{+}10, PK03]. Workshop [Pet03, PNPP12, PU04, PVTG07, PTZ03, Pet92, Pet02, PNB04, Pfi00, PMD90, PMD92, PMD94, PMDA96, PMD97, PMDK00, POS00, PG01, Pie05, Pin95, PT99, PB08b, PKSS02, PB97, PLG06, Pol99a, Pol01, PZH96, PC00, PBRB08, PBS08, PvS07, PT06a, PT11, PF07a, PWY08, PMS08, Pri03a, Pri03b, PQ05, Pri91, PM09, Pro96, PS12c, PS11b, PB92, QCDM06, QBY^{+}12, QV90, Rah04, RF10, RiN12, RPR08, RM09, RWA06, RSV07, RS99a, Ran93, RFZ03, RVY05, RS05, RT12, RTZ08, Ras01, Rau04a, RWY97, RTA95, RA00, RST09, Rei88, Rei89, RBF06, REF^{+}06, RMS^{+}05, Ria08, Ria09, RtTMP10, RtTMP11, RtTM12, RM04a, RMB10, RK09, RK12, RGW01, RABR03, RF99a, RSL01, Rob06, RR04b, RH01, RK02, RKW04, Rol97, RV02, RB10, RF12]. Workshop [RBSL06, RPZ97, RH98, RFBD09, RM04b, Rue93, SST00, Sab10, SBV^{+}11, SS08a, SW04, ŞS05a, ŞSW07, SSVW12, SFS11, SGSV10, SL11, SdSMO12, SS11b, SÇS12, SW09a, STP^{+}10, ST88, SH12a, SKR11, SHM^{+}10, SGG06, SMBQ07, Sch10a, Sch04a, STT11, ST08, SS01, Sch93a, SH12b, Sch93b, Sch03a, SKP10, SS94, Sch06, SE94, Sch01a, SW03a, SvS00, SS05b, SJDT08, SNRG11, SW03b, SWAF95, SM06, SE10, ST12b, SLH04, SLH05, SLZH07, SZ93, SZ92, Seh97, SHGB05, SHW09, Sen10, SSJR10, SG09, SOS96, SBC^{+}94, She01, She03, SS06b, SJ06, SKK08, SKPG12, SCG98, SBD03, SA06, Sif90, SRW98, SL99a, SLL01, Smi01, Smi05, SF96, Smi06, SKP00, ST07, SSM08, SPSV^{+}04, SOP12, SAUL11, SK97, SZ00, SK08, Som96, SMMG01, SPB00, SLPM96, SMK06]. Workshop [SH09, Squ09, SF08c, SS99, SMCM07, SS04, SvdMOP06, SSv94, Ste95, SW97, SP07, STWS07, SG07, SSKR02, SCKR03, SRT06, SHS10, ST01, SP12, SFV^{+}11, SGD09, SKB98, Str99, SZ11, SBH07, SBH09, SV01, SC07, SHUH08, SWSY11, SM09, Tah00, Tah01, TM09, TT95, TvdBI02, Tan11, TM99, TC95, TT10, TV94, TNN^{+}01, TD07, Thi10, TJ08, Tho11, TGH06, TS87, TBJ02, TQYX11, TC98, TG02, TT06b, TSK92, TST09, TZS00, Tro11, THR03, Tuc01, TtHVS06, TWSN07, UBM02, VvS99, VS09, Val06, VP96, VAR12, VWH06, VBH07, VEG^{+}11, Vau98, VY01, Vau05, Vck99, VA07, VDPD96, VC03, V^{+}06, VDD^{+}11, VSMS12, VBS^{+}10, VMB^{+}12, VT97, VZ99, VS05, VSTN06, Vos03, VAM06, VASP09, VKG12, VWY97, WF98, WS02, WK08, WR01, Wal84]. Workshop [WKP03, Wan98, WPRW06, WYSS10, WSN^{+}06, Waś96, Wei97, Wet92, WPM06, WPM07, WBD08, WMdLA10, WG11, WNE99, WB09, WGL96, WJP07, WR03, Win04, WNL05, WPH03, WKB04, WLS89, WHS01, Wol07, WY98, WCBD08, WJ95, WWC02, WVDS05, WD00, XW11, YD98, Yal10, YJ04, YJS^{+}06, YZC^{+}07, Yan11a, YLS^{+}12, YJKK06, YJKSW10, YC11, YSL12, YHV12, YSR^{+}06, YY09, Ytr06, YY01, ZSSM03, ZCCL10, ZGK^{+}07, ZCRL08, ZL96, ZL97, ZL98, ZS00, ZSM^{+}09, ZZZL12, ZG03, ZJL06, ZSO99, ZLY^{+}07a, ZJXC03, ZZTG07, ZT04, dBdRR89, dBdRR91, dB92, dBdRR93, dMB05, dMS06, dMGFS11, dS02, vB97, vDLJ11, vES10, vE91, vL88, vL89, vLS91, vL94b, vLS11, vdPW10, vdBK05, vdL98, vdL01, vdL02, vdL04, vdMBD08, vdMS11, vM94, vzGIK08, AYB95, AMW07, AZ96, A^{+}08]. workshop [ABLM05, AS06, Apt00, AGL06, Aus93, ADL09, BCG02, BL05a, Ban98, BDGM05, BCKK05, BBH^{+}05, BG91b, BE87, BB05a, BR06, BFS92, Bir07, BP01, BHR90, BL89, BC07, Bor91, Bor93, BGM94, BM98, Bre03, BJS93, BV05b, BS06c, BBC04, BTF05, BS06d, CC05, CNX05, CP06a, CHPW05, CR05a, CRW97, CFG96, CSV05, Coh92, CKMS05, Con00, Cow77, CKL05, DLT05, DSS91, Deh93, DLOS05, DY06b, DHX09, Der92, DJS91, DPP95, DBB^{+}04, DvV96, Dos93, Dov05, Dyc94, DHSH96, ELM05, Eng81, EHSH92, EW05, FS05, FPRAW04, FRA05a, FK87, Fei03, FR95a, FR96, FRS05, FFRS05, FMO05, FPMW06, Fri04, Fro02, FW92, FLMS05, FM91, FLS05, GMV05, GJ86, GS05, GLW88, Gra05, GHMT05, GT05, GAN05, GM05, Gue05, GNR94, GS94, HI93, HH04]. workshop [HWP01, HPR05, HW06b, IWC04, Jan93, JSZ95, JNS05b, Jon80, KJ88, Kar91, KKM04, KN91, KIRK05, KS05b, Kow01, KC84, Kra05, KBC05, Lam96, LM93, LS92, Lau93b, LD97, LG94, Lei95, LM97, LV05, LY05, LMT^{+}04, Mac06, MML97, Mai88, MMM01, MØ05a, MD04, MR91, MS00a, MPPJ^{+}05, MH05, MS99a, Mon96, MT05b, MD99, MSS05a, MBA05, MSY00, MZF94, MC05, Neh88, NOOS05, OGM05, OAdGBP00, OFK05, OPKR05, PT95, PTRH05, PVV05, PS98a, PFCS05, PSO05, Pet01, PGBK04, PN93, Pic86, Pit86, Pre95, PT06b, RMR05, Ral94, RM97a, RB91, RBS93, RKM06, RF99b, San02, Sch95, SSW92, SPZ90, SH91, Sch92, SM91, Sel95, SDH05, SA96, SC98, SKY06, SLF91, Sta82, SS06d, Ste94, SLP05, TWN92, TEB09, TMP04, TYNM05, TAS05, Vää99]. workshop [War94a, Wat95, WS96, WAR94b, WPM05, WF77, WMT96, YI91, ZSD03, ZLJ05, ZGT05, Zip92, dBdRR90, dVGR04, vP93, FMdGC99]. Workshops [ABG^{+}06, ADB^{+}12a, ADB^{+}12b, Ami98, AGT98, AK12, AAR05, AHSC11, ACHH12, BGG^{+}12, BBK12, BCG^{+}11, BDC12, BMO02, BPD06, BFT^{+}08, BB02, Bru06b, BkHJ^{+}04, BH06d, CAD^{+}02, CHB^{+}12, CPSA10, CGMP11, CVLL12, Ceb08, CAS^{+}09b, CW11, CWC^{+}07, CUDL02, Cha09, CWiM^{+}09, Che99a, CLZ^{+}09, CLLD09, CBS^{+}11, CSN14, Col07, CS12c, CvRVSN12, DGT10c, DDS12, DF10, DMB^{+}11, DFPV11, DGJ^{+}05, DHtM^{+}12, DB98b, DsSL^{+}07, DR09, DvEF06, DS11b, ED06a, EYX^{+}05, EIRW06, Eug09, FWZH06, FIAP01, FL09, FK12b, FCD^{+}04, FT94, FMC12a, FMC12b, FMC12c, GCFA12, GRB^{+}07, Gho10, Gie08, GCC^{+}02a, GKMN10, GHI^{+}06, GVT^{+}11, GDG11, HRK^{+}07, HK12, HZK08, HKI^{+}09, HNRW06, HM02, HP09, HYPY10b, HNPP07, HMBD09, IB07, IHX^{+}08, JSS08, JPXS04b, Kam99, Kap10, Kie12, KH11a, KH11b, KLT^{+}07, Küh07, Kut12a, Kut12b]. Workshops [LMSS07, LMT00, LAF^{+}10, LCY^{+}11, MRY^{+}11, MT^{+}03, MTH05, MTH06a, MTH06b, MTH07a, MTH07b, MTH08, MHD09, MDH10c, MDH11a, MBJ^{+}07, MCdS^{+}08, NMM10, NSM07, NVSB^{+}07, NMP^{+}04, OBS12, ORB^{+}09, OBM11, OD12b, PAV^{+}10, PJC^{+}12, PCA^{+}09, PVC^{+}99, RKM^{+}12, RTB02, RBsC^{+}06, SHN07, SINK08, ST11, SPÖ^{+}10, SPON08, SCD^{+}10, SPC^{+}08, SKK04, SBF08, SC07, TNA^{+}10, THX^{+}07, TDK^{+}10, VMvN12, VHM07, WJL^{+}12, WZH^{+}12, WZH^{+}07, WSTI07, WHG07, WvdH03, XYZU11, YCC^{+}08, Yao95, YKF^{+}06, YMY^{+}10, YYH^{+}12, ZSY^{+}06, dCdF^{+}08, dSOSR06, tHBP08, ALS^{+}05, Ari02, BM98, BBC04, Cha03a, CPGL05, Dam99b, DER94, Est95, Fro02, JP03, Kow01, KMTU03, LMT^{+}04, MMM01, MØ05a, M^{+}06, MD99, OYY03, PV04, Rol98, Rol99, Rol00, SLL^{+}06, WTZ^{+}04, vEHD05, MTC04]. workshops-ECDM [OYY03]. World [ADP09, AOSV08, APV09, AK99, AMM99, BLNG10, BdBK03, BCT02c, Bv97, BGKS04, BJNT06, CW94a, CM95, CD09, CMMR12, Che99a, CLS07b, CDL12, DFP^{+}12, DGR06, Fur95, FU96, GB99, GOZ05, HL05, IMWZ09, KLR03, LSST07, LMT00, LABM^{+}12, LCDT08, LDC^{+}09, MCA^{+}07, NRSSV05, OROD07, OPT01, OPP04, PM96, PTZ03, PBBY04, Rei99, RMSS12, dSCP11, SS06c, SBK01, SV01, VP96, VPK00, VROD08, WL04, WWD99a, WWD99b, ZYW08, AMU03, BLS02, HWP01, OTZ01, Kit98, MVCM10, NM97, Wie06, Cag00]. worlds [Heu98, Heu00, Nar01a]. Worldwide [MMT97, MKT98]. WORM [MT^{+}03]. WOSE [MTH05]. WQoSR [SPSV^{+}04]. WRAC [HRR^{+}06, THR03]. Wrac’h [HCDL09]. WRLA [Dur12, Ölv10]. Wroclaw [CKW11a, CKW11b, KCK08, NGHJ07, WDKW10a, WDKW10b, ACJT07, BSS11, KCG09, NBGA08, NKC09]. WRSM [MT^{+}03]. WS [BKZ05, BNZ06, BB11, BW09, CP12, DH08, LS10a, JPXS04b]. WS-FM [BKZ05, BNZ06, BB11, BW09, CP12, DH08, LS10a]. WSA [Cou93b]. WSCOBPM [BH06d]. WSKS [LCDT08, LDC^{+}09]. WSOC [DsSL^{+}07]. WSOM [LH11, PM09]. WSS [DH01]. WTCS [DKN12]. Wuhan [APR^{+}06, CHKL10, FWZH06, JWH06, JPXS04a, JPXS04b, JGXC04, KLZ07a, KLZ07b, KCYL08, LCY10, LWD05, MJYT06, WLO^{+}11, WJL^{+}12, XHXL08, XLHX08, YJMU06, YHZ09a, YHZ09b, YHZ09c]. Würzburg [EFW93, Pup99, SHW09]. WUsCAM [Bru06b]. Wusterhausen [Ehr91]. Wusterhausen/Dosse [Ehr91]. Wuxi [LFJI10a, LFJI10b, LJS^{+}10]. Wuyishan [X^{+}12a, X^{+}12b, XBM12]. WWCA [MMT97]. WWCA’98 [MKT98]. WWIC [BMM07, BCKT05, BCF06, HHLS08, KMMT12, MBVTY11, OKBMB10, vdBHOS09, LLMT04]. WWW [ZYBX08, ZLS^{+}06]. Wye [CH91].

X [ADP09, DP10a, DFN10, GBH11,

GTM10a, GGS11, KRP06, RR89, RJL^{+}08, LSST07, PSW^{+}09, PDAN08, Spa08]. X-MP [RR89]. X-MP/Model [RR89]. Xanthi [BCKT05, SBH^{+}04]. XHPC [M^{+}06]. XI [BGJ11, DLGJ10, GTM10b, HKT12, KHPS07, MBS05a, SCKR10a, SCKR10b, WG11, PS10, PBP09, SPT^{+}08, VROD08]. Xiamen [DWML05, SLB05, ZJXC03]. Xi’an [ADDL08, ACK^{+}07, BLW09, BC11b, CLZ06, DXZ07, DL95b, H^{+}05, JWbG^{+}06, JWG^{+}06, LZL11, LZhL06, WYL01, WJS^{+}06, XBS^{+}10, YZZ^{+}05, YLC^{+}10, ZPT^{+}06, ZiTM10a, ZiTM10b, ZiTM10c, ZZZL12, ZJL06, vHV08, LOS05, MXZ05, QOZ01, ZZO03]. XII [CCD^{+}12a, CCD^{+}12b, GTSS11, KPP08, LTÅ^{+}11, MSW07, PBG^{+}10, VSMS12, IMWZ09, PSS^{+}10b, Spa09]. XIII [FvdTDG12, HMS09, BLNG10, GT11b, PSC^{+}11, PBPdV11, SZS09]. XIIth [PS11a]. Xining [PYLW08]. XIth [SC09a]. XIV [GTM11, MRU12, PPdV12, PSS^{+}11, dSCP11, SD09]. XIX [CCMS11, JW05]. XML [AYBH^{+}06, BBB^{+}07a, BHRU09, BCH^{+}05, FLMS05, FLMK06, FLT07, FKLT08, GKT09, GKT10, GKST11, GKS12, LYBU10, NZ06, SBH^{+}04, Bel03, BMR^{+}04, BGS^{+}03, Bre03, CUDL02]. XML-Based [CUDL02]. XMLDM [BGG^{+}12, SPÖ^{+}10, WJL^{+}12, CUDL02]. XP [AMS06, BMH05, CDSS07, ZEL04, EB04, MS03a, MW03, WW02]. XP/Agile [ZEL04, MW03, WW02]. XPS [Pup99]. XPS-99 [Pup99]. XRA [SLL^{+}06]. XSDM [JP03]. XSym [AYBH^{+}06, BBB^{+}07a, BHRU09, BCH^{+}05, LYBU10, BMR^{+}04, Bel03]. XV [GTP12, PS12b, RMSS12, Spa11]. XVI [CMMR11, GT12]. XVII [AHJ07a, AHJ07b, AHJ07c]. XX [CMSA12, DL06a]. XXI [BA07]. XXII [Atl08]. XXIII [GV09]. XXIV [CPSA10, FJ10]. XXV [Li11b]. XXVI [CBCGA12].

Y12M [ZWS81]. Ya’akov [PT99]. Yamaguchi [ZSO99]. Yamanaka [JMO10]. Yaroslavl [AN97, Mal97b]. Ybbs [MPPU11]. Year [DP10b]. Years [GCSR07, GV08, LL06, LIBP07, AJS05, Wil01b]. Yellow [JYZ04, WZJX12]. Yerevan [DV07b]. Yogyakarta [FV10]. Yokohama [Ari02, AiNOW11, BKNS06, GM06b, IZ98, KJS01, Mat02, UMN^{+}08, Yam08]. York [AP10, ATO^{+}09, AT02c, AN07, BGKY08, BN04, BDDS05, BDGK00, BGHS05, CP98a, CDP^{+}06, Can08, Cas09a, CW09, CPPB06, DLT05, DGJ^{+}05, DV09, DIS11, DPP98, DCFY10, EKS00, FRA05a, FRS05, GW11b, HR06, HP99b, HV03, HMS09, Heb95, IKY05, KP09, Koz82, LOTY05, Lov82, MPP94, MAFS08a, MAFS08b, MA11, MBA05, NKN^{+}05, OGM05, Pac05, Per11a, Per11b, RMR05, Sch01a, SC00, TTM10, UDF06, VW05, WPM05, Yao82, YDKM06, Zuc03, dA09, dMS09, vEHD05]. Yorktown [Koz82, MMG^{+}01]. Young [DK90]. YRWS [CUDL02]. Ystad [HK11]. Yucatan [Coe02]. Yuri [BDR10]. Yvette [Bra99, CMT01].

Z [BHL90, BBBB08, DFG^{+}12, FGK^{+}10, TKHS05, Ber02b, BBKW03, BH95, BHT97, BFH98, BDGK00]. Zaborow [Sko85]. Zagreb [LHJ08a, LHJ08b, LHJ08c]. Zakopane [RM96, RSTZ04, RTZŻ06, RTZZ08, RST^{+}10b, RST^{+}10a, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, Win78]. Zalessky [Mal07, MT89]. Zandvoort [Kat92]. Zaragoza [BBP04, GTT04, GM08a, KMW04, KLP04, Tra04b, Val94]. Zarós [Thi10]. ZB [Ber02b, BBKW03, BDGK00, TKHS05]. Zdzisław [PSD^{+}07, PSM^{+}07]. Zealand [BMS^{+}08a, BK05a, BBKM11, BS06c, CCD04, DKN12, ET01a, HRK^{+}07, HZK08, JP08, Kie12, KKS11c, KKS11d, KKS11e, KKS11f, KPS01, KŽ04, KH11a, KH11b, KKC09a, KKC09b, MJR04, NHJ04a, NHJ04b, NHJ04c, PSST07, PS09c, SK08, WCK11, WZ08, ZGY04, Zhe02]. Zeist [EGO09]. Zelezna [Mat01]. Zermatt [MT09b]. Zhangjiajie [CLL^{+}09, LZ05, WZD11]. Zhengzhou [DSZ10, HGBF12, HGGG12, HGPH12, QIW07]. ZhenJiang [CL04]. Zhou [JLW07a]. Zhuhai [KY07a]. ZiF [WK08]. Zikhron [PT99]. Zikhron-Ya’akov [PT99]. Zohar [Der04]. Zuckerman [Gol11b]. ZUM [BH95, BHT97]. ZUM’98 [BFH98]. Zurich [AHS11a, AHS11b, AJ11, AE06, BF09, BV11, BM06, CV11, Dub09, EK07, FRvKS10, FLF^{+}08, GW11a, GM07b, HS06, HKNW07, HKV10, KP10b, KF12, LTT07, MJ07, MW08, Mic10, NG10, Pai09, RMB10, RLT^{+}08, SH09, SD10, BL06b, CFR92, Eng81, GKW02, GS91, Gun94, Gut94, GW00, JS97, KBD87, MT05b, Pla96, RGW01, RKM06, SS99, STTP02, SSKR02, VGP08, Vck99, WT96, WVD02, Zit01, vG02, KM00, Bur90, MN02b]. Zvika [Gol11b].

References

MISSING-1588

[???a]  ????, editor.  ????,

volume 1588 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=1588.

MISSING-2354

[???b]  ????, editor.  ????,

volume 2354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=

issue&issn=0302-9743&
volume=2354.

MISSING-2579

[???c]  ????, editor.  ????,

volume 2579 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=2579.

MISSING-2974

[???d]  ????, editor.  ????,

volume 2974 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=2974.

MISSING-3197

[???e]  ????, editor.  ????,

volume 3197 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3197.

MISSING-3227

[???f]  ????, editor.  ????,

volume 3227 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3227.

MISSING-4380

[???g]  ????, editor.  ????,

volume 4380 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=4380.

MISSING-4920

[???h]  ????, editor.  ????,

volume 4920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=4920.

MISSING-6103

[???i]  ????, editor.  ????,

volume 6103 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=6103.

MISSING-6126

[???j]  ????, editor.  ????,

volume 6126 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=6126.

MISSING-6127

[???k]  ????, editor.  ????,

volume 6127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=6127.

MISSING-6368

[???l]  ????, editor.  ????,

volume 6368 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=6368.

MISSING-6429

[???m]  ????, editor.  ????,

volume 6429 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=6429.

MISSING-6745

[???n]  ????, editor.  ????,

volume 6745 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=6745.

MISSING-6983

[???o]  ????, editor.  ????,

volume 6983 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=6983.

MISSING-7028

[???p]  ????, editor.  ????,

volume 7028 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7028.

MISSING-7061

[???q]  ????, editor.  ????,

volume 7061 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7061.

MISSING-7070

[???r]  ????, editor.  ????,

volume 7070 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7070.

MISSING-7123

[???s]  ????, editor.  ????,

volume 7123 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7123.

MISSING-7146

[???t]  ????, editor.  ????,

volume 7146 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7146.

MISSING-7171

[???u]  ????, editor.  ????,

volume 7171 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7171.

MISSING-7229

[???v]  ????, editor.  ????,

volume 7229 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7229.

MISSING-7263

[???w]  ????, editor.  ????,

volume 7263 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7263.

MISSING-7326

[???x]  ????, editor.  ????,

volume 7326 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7326.

MISSING-7403

[???y]  ????, editor.  ????,

volume 7403 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7403.

MISSING-7407

[???z]  ????, editor.  ????,

volume 7407 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7407.

MISSING-7420

[???-27]  ????, editor.  ????,

volume 7420 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7420.

MISSING-7423

[???-28]  ????, editor.  ????,

volume 7423 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7423.

MISSING-7475

[???-29]  ????, editor.  ????,

volume 7475 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7475.

MISSING-7480

[???-30]  ????, editor.  ????,

volume 7480 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7480.

MISSING-7500

[???-31]  ????, editor.  ????,

volume 7500 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7500.

MISSING-7536

[???-32]  ????, editor.  ????,

volume 7536 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7536.

MISSING-7537

[???-33]  ????, editor.  ????,

volume 7537 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7537.

MISSING-7540

[???-34]  ????, editor.  ????,

volume 7540 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7540.

MISSING-7542

[???-35]  ????, editor.  ????,

volume 7542 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7542.

MISSING-7544

[???-36]  ????, editor.  ????,

volume 7544 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7544.

MISSING-7545

[???-37]  ????, editor.  ????,

volume 7545 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7545.

MISSING-7565

[???-38]  ????, editor.  ????,

volume 7565 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7565.

MISSING-7566

[???-39]  ????, editor.  ????,

volume 7566 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7566.

MISSING-7567

[???-40]  ????, editor.  ????,

volume 7567 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7567.

MISSING-7581

[???-41]  ????, editor.  ????,

volume 7581 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7581.

MISSING-7582

[???-42]  ????, editor.  ????,

volume 7582 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7582.

MISSING-7586

[???-43]  ????, editor.  ????,

volume 7586 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7586.

MISSING-7587

[???-44]  ????, editor.  ????,

volume 7587 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7587.

MISSING-7588

[???-45]  ????, editor.  ????,

volume 7588 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7588.

MISSING-7591

[???-46]  ????, editor.  ????,

volume 7591 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7591.

MISSING-7593

[???-47]  ????, editor.  ????,

volume 7593 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7593.

MISSING-7605

[???-48]  ????, editor.  ????,

volume 7605 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7605.

MISSING-7606

[???-49]  ????, editor.  ????,

volume 7606 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7606.

MISSING-7607

[???-50]  ????, editor.  ????,

volume 7607 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7607.

MISSING-7627

[???-51]  ????, editor.  ????,

volume 7627 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7627.

MISSING-7629

[???-52]  ????, editor.  ????,

volume 7629 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7629.

MISSING-7630

[???-53]  ????, editor.  ????,

volume 7630 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7630.

MISSING-7640

[???-54]  ????, editor.  ????,

volume 7640 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7640.

MISSING-7641

[???-55]  ????, editor.  ????,

volume 7641 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7641.

MISSING-7642

[???-56]  ????, editor.  ????,

volume 7642 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7642.

MISSING-7643

[???-57]  ????, editor.  ????,

volume 7643 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7643.

MISSING-7644

[???-58]  ????, editor.  ????,

volume 7644 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7644.

MISSING-7647

[???-59]  ????, editor.  ????,

volume 7647 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7647.

MISSING-7648

[???-60]  ????, editor.  ????,

volume 7648 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7648.

MISSING-7649

[???-61]  ????, editor.  ????,

volume 7649 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7649.

MISSING-7650

[???-62]  ????, editor.  ????,

volume 7650 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7650.

MISSING-7651

[???-63]  ????, editor.  ????,

volume 7651 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7651.

MISSING-7652

[???-64]  ????, editor.  ????,

volume 7652 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7652.

MISSING-7656

[???-65]  ????, editor.  ????,

volume 7656 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7656.

MISSING-7657

[???-66]  ????, editor.  ????,

volume 7657 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7657.

MISSING-7660

[???-67]  ????, editor.  ????,

volume 7660 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7660.

MISSING-7662

[???-68]  ????, editor.  ????,

volume 7662 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7662.

MISSING-7669

[???-69]  ????, editor.  ????,

volume 7669 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7669.

MISSING-7676

[???-70]  ????, editor.  ????,

volume 7676 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7676.

MISSING-7680

[???-71]  ????, editor.  ????,

volume 7680 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7680.

MISSING-7681

[???-72]  ????, editor.  ????,

volume 7681 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7681.

MISSING-7684

[???-73]  ????, editor.  ????,

volume 7684 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7684.

MISSING-7685

[???-74]  ????, editor.  ????,

volume 7685 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7685.

MISSING-7686

[???-75]  ????, editor.  ????,

volume 7686 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7686.

MISSING-7687

[???-76]  ????, editor.  ????,

volume 7687 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7687.

MISSING-7688

[???-77]  ????, editor.  ????,

volume 7688 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7688.

MISSING-7689

[???-78]  ????, editor.  ????,

volume 7689 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7689.

MISSING-7691

[???-79]  ????, editor.  ????,

volume 7691 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7691.

MISSING-7692

[???-80]  ????, editor.  ????,

volume 7692 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7692.

MISSING-7693

[???-81]  ????, editor.  ????,

volume 7693 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=7693.

MISSING-7694

[???-82]  ????, editor.  ????,

volume 7694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7694.

MISSING-7695

[???-83]  ????, editor.  ????,

volume 7695 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7695.

MISSING-7696

[???-84]  ????, editor.  ????,

volume 7696 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, ???? CODEN LNCSD9. ISBN ???? ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=7696.

Aiello:2008:AMW

[A^{+}08] William Anthony Aiello et al., editors. Algorithms and models for the webgraph: fourth international workshop, WAW 2006, Banff, Canada, November 30–December 1, 2006: revised papers, volume 4936 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. ISBN 3-540-78808-5, 3-540-78807-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W39 2006. URL http:

//www.springerlink.
com/content/978-3-540-
78808-9.

Agon:2011:MCM

[AAA^{+}11] Carlos Agon, Moreno Andreatta, Gérard Assayag, Emmanuel Amiot, Jean Bresson, and John Mandereau, editors. Mathematics and Computation in Music: Third International Conference, MCM 2011, Paris, France, June 15–17, 2011. Proceedings, volume 6726 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21589-0 (print), 3-642-21590-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
21590-2.

Almeida:2006:HMT

[AAB^{+}06] Francisco Almeida, María J. Blesa Aguilera, Christian Blum, José Marcos Moreno Vega, Melquíades Pérez Pérez, Andrea Roli, and Michael Sampels, editors. Hybrid Metaheuristics: Third International Workshop, HM 2006 Gran Canaria, Spain, October 13–14, 2006 Proceedings, volume 4030 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46384-4 (print), 3-540-46385-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46385-6.

Antoniou:2007:RWT

[AAB^{+}07] Grigoris Antoniou, Uwe Aßmann, Cristina Baroglio, Stefan Decker, Nicola Henze, Paula-Lavinia Patranjan, and Robert Tolksdorf, editors. Reasoning Web: Third International Summer School 2007, Dresden, Germany, September 3–7, 2007, Tutorial Lectures, volume 4636 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74613-7 (print), 3-540-74615-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74615-7.

Aroyo:2010:SWRb

[AAH^{+}10a] Lora Aroyo, Grigoris Antoniou, Eero Hyvönen, Annette ten Teije, Heiner Stuckenschmidt, Liliana Cabral, and Tania Tudorache, editors. The Semantic Web: Research and Applications: 7th Extended Semantic Web Conference, ESWC 2010, Heraklion, Crete, Greece, May 30 — June 3, 2010, Proceedings, Part II, volume 6089 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13488-2 (print), 3-642-13489-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13489-0.

Aroyo:2010:SWRa

[AAH^{+}10b] Lora Aroyo, Grigoris Antoniou, Eero Hyvönen, Annette ten Teije, Heiner Stuckenschmidt, Liliana Cabral, and Tania Tudorache, editors. The Semantic Web: Research and Applications: 7th Extended Semantic Web Conference, ESWC 2010, Heraklion, Crete, Greece, May 30 –June 3, 2010, Proceedings, Part I, volume 6088 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13485-8 (print), 3-642-13486-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13486-9.

Azvine:2000:ISS

[AAN00] Benham Azvine, Nader Azarmi, and Detlef Nauck. Intelligent systems and soft computing: prospects, tools, and applications, volume 1804 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67837-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xvii + 357 pp. LCCN QA76.76.E95 .I5654 2000.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1804.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67837-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1804.

Albers:2009:EAE

[AAN09] Susanne Albers, Helmut Alt, and Stefan Näher, editors. Efficient Algorithms: Essays Dedicated to Kurt Mehlhorn on the Occasion of His 60th Birthday, volume 5760 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03455-1 (print), 3-642-03456-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03456-5.

Assman:2005:MDA

[AAR05] Uwe Aßman, Mehmet Aksit, and Arend Rensink, editors. Model Driven Architecture: European MDA Workshops: Foundations and Applications, MDAFA 2003 and MDAFA 2004, Twente, The Netherlands, June 26–27, 2003 and Linköping, Sweden, June 10–11, 2004. Revised Selected Papers, volume 3599 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28240-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .M33 2003. URL http:

//www.springerlink.
com/content/978-3-540-
28240-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3599.

Aagesen:2004:ICS

[AAW04] Finn Arve Aagesen, Chutiporn Anutariya, and Vilas Wuwongse, editors. Intelligence in Communication Systems IFIP International Conference, INTELL-COMM 2004, Bangkok, Thailand, November 23– 26, 2004: Proceedings, volume 3283 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23893-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3283.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23893-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3283;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b103336.

Ausiello:1978:ALP

[AB78] G. Ausiello and C. Bohm, editors. Automata, languages and programming: fifth colloquium, Udine, Italy, July 17– 21, 1978, volume 62 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1978. CO-DEN LNCSD9. ISBN 0-387-08860-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.62.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0062.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        08860-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=62.

Astesiano:1981:TAP

[AB81] E. Astesiano and C. Bohm, editors. Trees in algebra and programming: 6th colloquium, Genoa, March 5–7, 1981: proceedings, volume 112 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10828-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166.2 .C3 1981. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0112.htm;
 http://www.springerlink.
com/content/978-0-387-
10828-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=112.

Adam:1993:ADS

[AB93] Nabil R. Adam and Bharat K. Bhargava. Advanced database systems, volume 759 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57507-3 (Berlin), 0-387-57507-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 451 pp. LCCN QA76.9.D3 A3433 1993. DM86.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0759.htm;
 http://www.springerlink.
com/content/978-3-540-
57507-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=759.

Azema:1997:ATP

[AB97] Pierre Azema and Gianfranco Balbo, editors. Application and theory of Petri nets 1997: 18th international conference, ICATPN’97, Toulouse, France, June 23–27, 1997: proceedings, volume 1248 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63139-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A66 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1248.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63139-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1248.

Ashley:2003:CBR

[AB03] K. D. Ashley and Derek G. Bridge, editors. Casebased reasoning research and development: 5th International Conference on Case-Based Reasoning, ICCBR 2003, Trondheim, Norway, June 23–26, 2003: Proceedings, volume 2689 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40433-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 C376 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2689.htm;
 http://www.springerlink.
com/content/978-3-540-
40433-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2689.

Antoniou:2004:RRM

[AB04] Grigoris Antoniou and Harold Boley, editors. Rules and Rule Markup Languages for the Semantic Web: Third International Workshop, RULEML 2004, Hiroshima, Japan, November 8, 2004, Proceedings, volume 3323 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23842-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3323.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23842-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3323;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b102922.

Asarin:2006:FMA

[AB06] Eugene Asarin and Patricia Bouyer, editors. Formal Modeling and Analysis of Timed Systems: 4th International Conference, FORMATS 2006, Paris, France, September 25–27, 2006. Proceedings, volume 4202 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45026-2 (print), 3-540-45031-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        45031-3.

Abdalla:2010:PCL

[AB10] Michel Abdalla and Paulo S. L. M. Barreto, editors. Progress in Cryptology — LAT-INCRYPT 2010: First International Conference on Cryptology and Information Security in Latin America, Puebla, Mexico, August 8–11, 2010, proceedings, volume 6212 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14711-9 (print), 3-642-14712-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14712-8.

Achterberg:2011:IAT

[AB11] Tobias Achterberg and J. Christopher Beck, editors. Integration of AI and OR Techniques in Constraint Programming for Combinatorial Optimization Problems: 8th International Conference, CPAIOR 2011, Berlin, Germany, May 23–27, 2011. Proceedings, volume 6697 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21310-3 (print), 3-642-21311-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21311-3.

Ailamaki:2012:SSD

[AB12] Anastasia Ailamaki and Shawn Bowers, editors. Scientific and Statistical Database Management: 24th International Conference, SSDBM 2012, Chania, Crete, Greece, June 25–27, 2012. Proceedings, volume 7338 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31234-9 (print), 3-642-31235-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31235-9.

Althoff:1999:CBR

[ABB99]  Klaus-Dieter  Althoff,

R. Bergmann, and Karl Branting, editors. Casebased reasoning research and development: Third International Conference on Case-Based Reasoning, ICCBR-99, Seeon Monastery, Germany, July 27–30, 1999: Proceedings, volume 1650 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66237-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I55 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1650.htm;
http://www.springerlink.
com/content/978-3-540-
66237-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1650.

Allbeck:2010:IVA

[ABB^{+}10a] Jan Allbeck, Norman Badler, Timothy Bickmore, Catherine Pelachaud, and Alla Safonova, editors. Intelligent Virtual Agents: 10th International Conference, IVA 2010, Philadelphia, PA, USA, September 20– 22, 2010. Proceedings, volume 6356 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15891-9 (print), 3-642-15892-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15892-6.

Alpcan:2010:DGT

[ABB10b] Tansu Alpcan, Levente Buttyán, and John S.

Baras, editors. Decision and Game Theory for Security: First International Conference, GameSec 2010, Berlin, Germany, November 22– 23, 2010. Proceedings, volume 6442 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17196-6 (print), 3-642-17197-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17197-0.

Aggarwal:2011:CIA

[ABB^{+}11] Jake K. Aggarwal, Reneta P. Barneva, Valentin E. Brimkov, Kostadin N. Koroutchev, and Elka R. Korutcheva, editors. Combinatorial Image Analysis: 14th International Workshop, IWCIA 2011, Madrid, Spain, May 23– 25, 2011. Proceedings, volume 6636 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21072-4 (print), 3-642-21073-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21073-0.

Abello:2012:MDE

[ABB12] Alberto Abelló, Ladjel Bellatreche, and Boualem Benatallah, editors. Model and Data Engineering: 2nd International Conference, MEDI 2012, Poitiers, France, October 3–5, 2012. Proceedings, volume 7602 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33608-6 (print), 3-642-33609-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33609-6.

Aldini:2010:CPE

[ABBC10] Alessandro Aldini, Marco Bernardo, Luciano Bononi, and Vittorio Cortellessa, editors. Computer Performance Engineering: 7th European Performance Engineering Workshop, EPEW 2010, Bertinoro, Italy, September 23–24, 2010. Proceedings, volume 6342 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15783-1 (print), 3-642-15784-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15784-4.

Al-Begain:2011:ASM

[ABBFM11] Khalid Al-Begain, Simonetta Balsamo, Dieter Fiems, and Andrea Marin, editors. Analytical and Stochastic Modeling Techniques and Applications: 18th International Conference, ASMTA 2011, Venice, Italy, June 20–22, 2011. Proceedings, volume 6751 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21712-5 (print), 3-642-21713-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21713-5.

Ahlswede:2006:GTI

[ABC^{+}06] Rudolf Ahlswede, Lars Bäumer, Ning Cai, Harout Aydinian, Vladimir Blinovsky, Christian Deppe, and Haik Mashurian, editors. General Theory of Information Transfer and Combinatorics, volume 4123 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46244-9 (print), 3-540-46245-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46245-3.

Akoka:1999:CME

[ABCWM99] Jacky Akoka, Mokrane Bouzeghoub, Isabell Comyn-Wattiau, and Elisabeth Métais, editors. Conceptual modeling — ER’99: 18th International Conference on Conceptual Modeling, Paris, France, November 15–18, 1999: Proceedings, volume 1728 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66686-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1728.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1728.
htm;   http://www.springerlink.
com/content/978-3-540-
66686-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1728.

Amestoy:1999:EPP

[ABD^{+}99] Patrick Amestoy, Philippe Berger, Michel Daydé, Iain Duff, Valérie Frayssé, Luc Giraud, and Daniel Ruiz, editors. Euro-Par’99 parallel processing: 5th International Euro-Par Conference, Toulouse, France, August 31–September 3, 1999: Proceedings, volume 1685 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66443-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I553 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1685.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66443-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1685.

Asperti:2003:MKM

[ABD03] Andrea Asperti, Bruno Buchberger, and James Harold Davenport, editors. Mathematical Knowledge Management: Second International Conference, MKM 2003 Bertinoro, Italy, February 16–18, 2003, volume 2594 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00568-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .I565 2003. 65.50 SFR; 40.66 EUR. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2594.htm;
 http://www.springerlink.
com/content/978-3-540-
00568-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2594.

AlvarezLopez:2004:KEL

[ABD04] Jesús A. Alvarez López, Emilio Benfenati, and Werner Dubitzky, editors. Knowledge exploration in life science informatics: international symposium KELSI 2004, Milan, Italy, November 25–26, 2004, proceedings, volume 3303 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23927-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-

                                        23927-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3303;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b103729.

Armando:2008:ARI

[ABD08] Alessandro Armando, Peter Baumgartner, and Gilles Dowek, editors. Automated Reasoning: 4th International Joint Conference, IJCAR 2008 Sydney, Australia, August 12–15, 2008 Proceedings, volume 5195 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-71069-8 (print), 3-540-71070-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71070-7.

Abran:2009:SPP

[ABD^{+}09] Alain Abran, René Braungarten, Reiner R. Dumke, Juan J. Cuadrado-Gallego, and Jacob Brunekreef, editors. Software Process and Product Measurement: International Conferences IWSM 2009 and Mensura 2009 Amsterdam, The Netherlands, November 4–6, 2009. Proceedings, volume 5891 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05414-5 (print), 3-642-05415-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
05415-0.

Anastasi:2012:MTN

[ABD^{+}12] Giuseppe Anastasi, Emilio Bellini, Elisabetta Di Nitto, Carlo Ghezzi, Letizia Tanca, and Eugenio Zimeo, editors. Methodologies and Technologies for Networked Enterprises: ArtDeco: Adaptive Infrastructures for Decentralised Organisations, volume 7200 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31738-3 (print), 3-642-31739-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31739-2.

Aloni:2010:LLM

[ABdJS10] Maria Aloni, Harald Bastiaanse, Tikitu de Jager, and Katrin Schulz, editors. Logic, Language and Meaning: 17th Amsterdam Colloquium, Amsterdam, The Netherlands, December 16–18, 2009, Revised Selected Papers, volume 6042 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14286-9 (print), 3-642-14287-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14287-1.

Aldini:2010:FMQ

[ABDW10] Alessandro Aldini, Marco Bernardo, Alessandra Di Pierro, and Herbert Wiklicky, editors. Formal Methods for Quantitative Aspects of Programming Languages: 10th International School on Formal Methods for the Design of Computer, Communication and Software Systems, SFM 2010, Bertinoro, Italy, June 21– 26, 2010, Advanced Lectures, volume 6154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13677-X (print), 3-642-13678-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13678-8.

Abe:2006:TCC

[Abe06] Masayuki Abe, editor.

Topics in Cryptology — CT-RSA 2007: The Cryptographers’ Track at the RSA Conference 2007, San Francisco, CA, USA, February 5–9, 2007. Proceedings, volume 4377 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69327-0 (print), 3-540-69328-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69328-4.

Abe:2010:ACA

[Abe10] Masayuki Abe, editor.

Advances in Cryptology — ASIACRYPT 2010: 16th International Conference on the Theory and Application of Cryptology and Information Security, Singapore, December 5–9, 2010. Proceedings, volume 6477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17372-1 (print), 3-642-17373-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17373-8.

AAgotnes:2012:DLC

[ÅBE12] Thomas Ågotnes, Jan Broersen, and Dag Elgesem, editors. Deontic Logic in Computer Science: 11th International Conference, DEON 2012, Bergen, Norway, July 16–18, 2012. Proceedings, volume 7393 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31569-0 (print), 3-642-31570-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31570-1.

Airoldi:2007:SNA

[ABF^{+}07] Edoardo Airoldi, David M.

Blei, Stephen E. Fienberg, Anna Goldenberg, Eric P. Xing, and Alice X. Zheng, editors. Statistical Network Analysis: Models, Issues, and New Directions: ICML 2006 Workshop on Statistical Network Analysis, Pittsburgh, PA, USA, June 29, 2006, Revised Selected Papers, volume 4503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73132-6 (print), 3-540-73133-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73133-7.

Al-Begain:2009:ASM

[ABFH09] Khalid Al-Begain, Dieter Fiems, and Gábor Horváth, editors. Analytical and Stochastic Modeling Techniques and Applications: 16th International Conference, ASMTA 2009, Madrid, Spain, June 9–12, 2009. Proceedings, volume 5513 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02204-9 (print), 3-642-02205-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02205-0.

Al-Begain:2010:ASM

[ABFK10] Khalid Al-Begain, Dieter Fiems, and William J. Knottenbelt, editors. Analytical and Stochastic Modeling Techniques and Applications: 17th International Conference, ASMTA 2010, Cardiff, UK, June 14–16, 2010. Proceedings, volume 6148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13567-6 (print), 3-642-13568-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13568-2.

Azevedo:2007:RAC

[ABFR07] Francisco Azevedo, Pedro Barahona, François Fages, and Francesca Rossi, editors. Recent Advances in Constraints: 11th Annual ERCIM International Workshop on Constraint Solving and Contraint Logic Programming, CSCLP 2006, Caparica, Portugal, June 26–28, 2006, Revised Selected and Invited Papers, volume 4651 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73816-9 (print), 3-540-73817-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73817-6.

Al-Begain:2012:ASM

[ABFV12] Khalid Al-Begain, Dieter Fiems, and Jean-Marc Vincent, editors. Analytical and Stochastic Modeling Techniques and Applications: 19th International Conference, ASMTA 2012, Grenoble, France, June 4–6, 2012. Proceedings, volume 7314 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30781-7 (print), 3-642-30782-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30782-9.

Apers:1996:ADT

[ABG96] P. M. G. (Petrus Maria Gerardus) Apers, M. (Mokrane) Bouzeghoub, and G. (Georges) Gardarin, editors. Advances in database technology — EDBT ’96: 5th International Conference on Extending Database Technology, Avignon, France, March 25–29, 1996: proceedings, volume 1057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61057-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55844 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1057.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61057-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1057.

Ackermann:2006:SWM

[ABG^{+}06] Markus Ackermann, Bettina Berendt, Marko Grobelnik, Andreas Hotho, Dunja Mladenič, Giovanni Semeraro, Myra Spiliopoulou, Gerd Stumme, Vojtěch Svátek, and Maarten van Someren, editors. Semantics, Web and Mining: Joint International Workshops, EWMF 2005 and KDO 2005, Porto, Portugal, October 3–7, 2005, Revised Selected Papers, volume 4289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47697-0 (print), 3-540-47698-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47698-6.

Aguirre:2009:MAA

[ABG09a] Arturo Hernández Aguirre, Raúl Monroy Borja, and Carlos Alberto Reyes Garciá, editors. MI-CAI 2009: Advances in Artificial Intelligence: 8th Mexican International Conference on Artificial Intelligence, Guanajuato, México, November 9–13, 2009. Proceedings, volume 5845 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05257-6 (print), 3-642-05258-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05258-3.

Aldini:2009:FSA

[ABG09b] Alessandro Aldini, Gilles Barthe, and Roberto Gorrieri, editors. Foundations of Security Analysis and Design V: FOSAD 2007/2008/2009 Tutorial Lectures, volume 5705 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03828-X (print), 3-642-03829-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03829-7.

Archibald:2009:ILC

[ABGL09] Margaret Archibald, Vasco Brattka, Valentin Goranko, and Benedikt Löwe, editors. Infinity in Logic and Computation: International Conference, ILC 2007, Cape Town, South Africa, November 3–5, 2007, Revised Selected Papers, volume 5489 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03091-2 (print), 3-642-03092-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03092-5.

Atkinson:2005:CBS

[ABGP05] Colin Atkinson, Christian Bunse, Hans-Gerhard Gross, and Christian Peper, editors. Componentbased software development for embedded systems: an overview of current research trends, volume 3778 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30644-7. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 344 pp. LCCN QA76.76.C66 C647 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30644-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3778.

Ando:2010:SMP

[ABH^{+}10] Noriaki Ando, Stephen Balakirsky, Thomas Hemker, Monica Reggiani, and Oskar von Stryk, editors. Simulation, Modeling, and Programming for Autonomous Robots: Second International Conference, SIM-PAR 2010, Darmstadt, Germany, November 15– 18, 2010. Proceedings, volume 6472 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17318-7 (print), 3-642-17319-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17319-6.

Adam:2012:MMC

[ABH12] Gheorghe Adam, Ján Buša, and Michal Hnatič, editors. Mathematical Modeling and Computational Science: International Conference, MMCP 2011, Stará Lesná, Slovakia, July 4– 8, 2011, Revised Selected Papers, volume 7125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28211-3 (print), 3-642-28212-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28212-6.

Al-Begain:2008:ASM

[ABHT08] Khalid Al-Begain, Armin Heindl, and Miklós Telek, editors. Analytical and Stochastic Modeling Techniques and Applications: 15th International Conference, ASMTA 2008 Nicosia, Cyprus, June 4–6, 2008 Proceedings, volume 5055 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68980-X (print), 3-540-68982-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68982-9.

Andersson:2002:NLP

[ABJ02] Birger Andersson, Maria Bergholtz, and Paul Johannesson, editors. Natural Language Processing and Information Systems: 6th International Conference on Applications of Natural Language to Information Systems, NLDB 2002, Stockholm, Sweden, June 27–28, 2002. Revised Papers, volume 2553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00307-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 I558 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2553.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00307-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2553.

Akiyama:2005:CGG

[ABK05] Jin Akiyama, Edy Tri Baskoro, and Mikio Kano, editors. Combinatorial geometry and graph theory: Indonesia-Japan joint conference, IJCCGGT 2003, Bandung, Indonesia, September 13–16, 2003. Revised selected papers, volume 3330 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24401-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA167 .I53 2003.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        24401-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3330;
                                         http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b105263.

Agosti:2009:RAT

[ABK^{+}09] Maristella Agosti, José Borbinha, Sarantos Kapidakis, Christos Papatheodorou, and Giannis Tsakonas, editors. Research and Advanced Technology for Digital Libraries: 13th European Conference, ECDL 2009, Corfu, Greece, September 27 — October 2, 2009. Proceedings, volume 5714 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04345-3 (print), 3-642-04346-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04346-8.

Andreev:2012:ITS

[ABK12] Sergey Andreev, Sergey Balandin, and Yevgeni Koucheryavy, editors. Internet of Things, Smart Spaces, and Next Generation Networking: 12th International Conference, NEW2AN 2012, and 5th Conference, ruS-MART 2012, St. Petersburg, Russia, August 27–29, 2012. Proceedings, volume 7469 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32685-4 (print), 3-642-32686-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32686-8.

Akiyama:2011:CGG

[ABKT11] Jin Akiyama, Jiang Bo, Mikio Kano, and Xuehou Tan, editors. Computational Geometry, Graphs and Applications: 9th International Conference, CGGA 2010, Dalian, China, November 3–6, 2010, Revised Selected Papers, volume 7033 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24982-5 (print), 3-642-24983-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24983-9.

Abrial:1996:FMI

[ABL96] Jean-Raymond Abrial, E. (Egon) Borger, and Hans Langmaack, editors. Formal methods for industrial applications: specifying and programming the steam boiler control, volume 1165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61929-1. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 509 pp. LCCN TJ288.F59 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1165.htm;
 http://www.springerlink.
com/content/978-3-540-
61929-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1165.

AjmoneMarsan:2005:QSM

[ABLM05] Marco Ajmone Marsan, Giuseppe Bianchi, Marco Listanti, and Michela Meo, editors. Quality of service in multiservice IP networks: third international workshop, QOS-IP 2005, Catania, Italy, February 2–4, 2004 proceedings, volume 3375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24557-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN TK5103.75 .Q67 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24557-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3375;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105169.

Anderson:2002:CSR

[ABM02] Stuart Anderson, Sandro Bologna, and Felici Massimo, editors. Computer safety, reliability, and security: 21st International Conference, SAFECOMP 2002, Catania, Italy, September 10–13, 2002: Proceedings, volume 2434 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44157-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 S23 2002; TJ212.2 .I545 2002. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2434.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        44157-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2434.

Ardissono:2005:UMI

[ABM05] Liliana Ardissono, Paul Brna, and Antonija Mitrovic, editors. User Modeling 2005: 10th International Conference, UM 2005, Edinburgh, Scotland, UK, July 24– 29, 2005. Proceedings, volume 3538 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27885-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I599 2005.

URL http://www.springerlink.
com/content/978-3-540-
27885-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3538.

Althoff:2008:ACB

[ABMH08]  Klaus-Dieter  Althoff,

Ralph Bergmann, Mirjam Minor, and Alexandre Hanft, editors. Advances in Case-Based Reasoning: 9th European Conference, EC-CBR 2008, Trier, Germany, September 1– 4, 2008. Proceedings, volume 5239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85501-7 (print), 3-540-85502-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85502-6.

Abrahamsson:2007:SPI

[ABMM07]  Pekka  Abrahamsson,

Nathan Baddoo, Tiziana Margaria, and Richard Messnarz, editors. Software Process Improvement: 14th European Conference, EuroSPI 2007, Potsdam, Germany, September 26– 28, 2007. Proceedings, volume 4764 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74765-6 (print), 3-540-75381-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75381-0.

Alferes:2006:PPS

[ABMS06] Jóse Júlio Alferes, James Bailey, Wolfgang May, and Uta Schwertel, editors. Principles and Practice of Semantic Web Reasoning: 4th International Workshop, PPSWR 2006, Budva, Montenegro, June 10–11, 2006, Revised Selected Papers, volume 4187 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39586-5 (print), 3-540-39587-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
39587-4.

Abramsky:2001:TLC

[Abr01] Samson Abramsky, editor. Typed lambda calculi and applications: 5th International Conference, TLCA 2001, Krakow, Poland, May 2–5, 2001: Proceedings, volume 2044 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41960-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.5 .I58 2001; Call number not reported. Internet. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2044.htm;
 http://www.springerlink.
com/content/978-3-540-
41960-0;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2044.

Abramowicz:2007:BIS

[Abr07] Witold Abramowicz, editor. Business Information Systems: 10th International Conference, BIS 2007, Poznan, Poland, April 25–27, 2007. Proceedings, volume 4439 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72034-0 (print), 3-540-72035-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72035-5.

Altmann:2009:GEB

[ABR09] Jörn Altmann, Rajkumar Buyya, and Omer F. Rana, editors. Grid Economics and Business Models: 6th International Workshop, GECON 2009, Delft, The Netherlands, August 24, 2009. Proceedings, volume 5745 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03863-8 (print), 3-642-03864-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03864-8.

Ayguade:2006:LCP

[ABRS06] Eduard Ayguadé, Gerald Baumgartner, J. Ramanujam, and P. Sadayappan, editors. Languages and Compilers for Parallel Computing: 18th International Workshop, LCPC 2005, Hawthorne, NY, USA, October 20–22, 2005, Revised Selected Papers, volume 4339 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69329-7 (print), 3-540-69330-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69330-7.

Astley:2006:DMI

[ABRZ06] Susan M. Astley, Michael Brady, Chris Rose, and Reyer Zwiggelaar, editors. Digital Mammography: 8th International Workshop, IWDM 2006, Manchester, UK, June 18–21, 2006. Proceedings, volume 4046 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35625-8 (print), 3-540-35627-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35627-1.

Andersen:1991:AIS

[ABS91] R. Andersen, Janis Bubenko, and Arne Sølvberg, editors. Advanced information systems engineering: Third International Conference CAiSE ’91, Trondheim, Norway, May 13–15, 1991: proceedings, volume 498 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54059-8 (Berlin), 0-387-54059-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .N67 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0498.htm;
 http://www.springerlink.
com/content/978-3-540-
54059-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=498.

Abowd:2001:UUC

[ABS01] Gregory D. Abowd, Barry Brumitt, and Steven Shafer, editors. Ubicomp 2001: ubiquitous computing: Third International Coference, Atlanta, Georgia, USA, September 30–October 2, 2001: Proceedings, volume 2201 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42614-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2201. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2201.htm;
http://www.springerlink.
com/content/978-3-540-
42614-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2201.

Asperti:2004:MKM

[ABT04] Andrea Asperti, Grzegorz Bancerek, and Andrzej Trybulec, editors. Mathematical Knowledge Management: Third International Conference, MKM 2004, Bialowieza, Poland, September 19– 21, 2004, Proceedings, volume 3119 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23029-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3119.htm;
 http://www.springerlink.
com/content/978-3-540-
23029-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3119;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100478.

Assmann:2010:RWS

[ABW10] Uwe Aßmann, Andreas Bartho, and Christian Wende, editors. Reasoning Web. Semantic Technologies for Software Engineering: 6th International Summer School 2010, Dresden, Germany, August 30

— September 3, 2010. Tutorial Lectures, volume 6325 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15542-1 (print), 3-642-15543-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15543-7.

Altman:2009:NCO

[AC09] Eitan Altman and Augustin Chaintreau, editors. Network Control and Optimization: Second Euro-NF Workshop, NET-COOP 2008 Paris, France, September 8–10, 2008. Revised Selected Papers, volume 5425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00392-3 (print), 3-642-00393-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00393-6.

Amati:2011:AIR

[AC11] Giambattista Amati and Fabio Crestani, editors. Advances in Information Retrieval Theory: Third International Conference, ICTIR 2011, Bertinoro, Italy, September 12–14, 2011. Proceedings, volume 6931 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23317-1 (print), 3-642-23318-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23318-0.

Artikis:2012:LPN

[ACÇ^{+}12] Alexander Artikis, Robert Craven, Nihan Kesim Çiçekli, Babak Sadighi, and Kostas Stathis, editors. Logic Programs, Norms and Action: Essays in Honor of Marek J. Sergot on the Occasion of His 60th Birthday, volume 7360 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29413-8 (print), 3-642-29414-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29414-3.

Akl:2007:UCI

[ACD^{+}07] Selim G. Akl, Cristian S. Calude, Michael J. Dinneen, Grzegorz Rozenberg, and H. Todd Wareham, editors. Unconventional Computation: 6th International Conference, UC 2007, Kingston, Canada, August 13–17, 2007. Proceedings, volume 4618 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73553-4 (print), 3-540-73554-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73554-0.

Autexier:2010:ICM

[ACD^{+}10] Serge Autexier, Jacques Calmet, David Delahaye, Patrick D. F. Ion, Laurence Rideau, Renaud Rioboo, and Alan P. Sexton, editors. Intelligent Computer Mathematics: 10th International Conference, AISC 2010, 17th Symposium, Calculemus 2010, and 9th International Conference, MKM 2010, Paris, France, July 5–10, 2010. Proceedings, volume 6167 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14127-7 (print), 3-642-14128-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14128-7.

Alvarez:2012:FIF

[ÁCD^{+}12] Federico Álvarez, Frances Cleary, Petros Daras, John Domingue, Alex Galis, Ana Garcia, Anastasius Gavras, Stamatis Karnouskos, Srdjan Krco, Man-Sze Li, Volkmar Lotz, Henning Müller, Elio Salvadori, Anne-Marie Sassen, Hans Schaffers, Burkhard Stiller, Georgios Tselentis, Petra Turkama, and Theodore Zahariadis, editors. The Future Internet: Future Internet Assembly 2012: From Promises to Reality, volume 7281 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30240-8 (print), 3-642-30241-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30241-1.

Aarts:2008:AIE

[ACdR^{+}08] Emile Aarts, James L. Crowley, Boris de Ruyter, Heinz Gerhäuser, Alexander Pflaum, Janina Schmidt, and Reiner Wichert, editors. Ambient Intelligence: European Conference, AmI 2008, Nuremberg, Germany, November 19– 22, 2008. Proceedings, volume 5355 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89616-3 (print), 3-540-89617-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89617-3.

Cesana:2006:WSN

[aCF06] atteo Cesana and Luigi Fratta, editors. Wireless Systems and Network Architectures in Next Generation Internet: Second International Workshop of the EURO-NGI Network of Excellence, Villa Vigoni, Italy, July 13– 15, 2005, revised selected papers, volume 3883 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-34025-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34025-6;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3883.

Aguzzoli:2007:APT

[ACG^{+}07] Stefano Aguzzoli, Agata Ciabattoni, Brunella Gerla, Corrado Manara, and Vincenzo Marra, editors. Algebraic and Proof-theoretic Aspects of Non-classical Logics: Papers in Honor of Daniele Mundici on the Occasion of His 60th birthday, volume 4460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75938-7 (print), 3-540-75939-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75939-3.

Anderson:2000:TAD

[ACH00] Michael Anderson, Peter Cheng, and Volker Haarslev, editors. Theory and Application of Diagrams: First International Conference, Diagrams 2000, Edinburgh, Scotland, UK, September 1–3, 2000. Proceedings, volume 1889 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67915-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA90 .D53 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1889.
htm;   http://www.springerlink.
com/content/978-3-540-
67915-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1889.

Atzmueller:2012:MMU

[ACHH12] Martin Atzmueller, Alvin Chin, Denis Helic, and Andreas Hotho, editors. Modeling and Mining Ubiquitous Social Media: International Workshops MSM 2011, Boston, MA, USA, October 9, 2011, and MUSE 2011, Athens, Greece, September 5, 2011, Revised Selected Papers, volume 7472 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33683-3 (print), 3-642-33684-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33684-3.

Atzeni:2008:ADI

[ACJ08] Paolo Atzeni, Albertas Caplinskas, and Hannu Jaakkola, editors. Advances in Databases and Information Systems: 12th East European Conference, ADBIS 2008, Pori, Finland, September 5–9, 2008. Proceedings, volume 5207 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85712-5 (print), 3-540-85713-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85713-6.

Alpuente:2009:FMI

[ACJ09] María Alpuente, Byron Cook, and Christophe Joubert, editors. Formal Methods for Industrial Critical Systems: 14th International Workshop, FMICS 2009, Eindhoven, The Netherlands, November 2–3, 2009. Proceedings, volume 5825 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04569-3 (print), 3-642-04570-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04570-7.

Arge:2007:ALP

[ACJT07] L. (Lars) Arge, Christian Cachin, Tomasz Jurdzinski, and Andrzej Tarlecki, editors. Automata, languages and programming: 34th international colloquium, ICALP 2007, Wrocław, Poland, July 9–13, 2007: proceedings, volume 4596 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-73419-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I55 2007. URL

http://www.springerlink.
com/content/978-3-540-
73419-2;   http://www.
springerlink.com/openurl.
asp?genre=book\%26isbn=
978-3-540-73419-2;
http://www.springerlink.
com/openurl.asp?genre=
issue\%26issn=0302-
9743\%26volume=4596.

Ambler:2000:SMI

[ACK00] Anthony Ambler, Seraphin B.

Calo, and Gautam Kar, editors. Services Management in Intelligent Networks: 11th IFIP/ IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2000, Austin, TX, USA, December 4–6, 2000. Proceedings, volume 1960 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41427-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I334 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1960.htm;
 http://www.springerlink.
com/content/978-3-540-
41427-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1960.

Akiyama:2007:DGC

[ACK^{+}07] Jin Akiyama, William Y. C. Chen, Mikio Kano, Xueliang Li, and Qinglin Yu, editors. Discrete Geometry, Combinatorics and Graph Theory: 7th China-Japan Conference, CJCDGCGT 2005, Tianjin, China, November 18–20, 2005, Xi’an, China, November 22–24, 2005, Revised Selected Papers, volume 4381 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70665-8 (print), 3-540-70666-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70666-3.

Andreasen:1998:FQA

[ACL98] Troels Andreasen, Henning Christiansen, and Henrik Legind Larsen, editors. Flexible query answering systems: Third International Conference, FQAS ’98, Roskilde, Denmark, May 13– 15, 1998: Proceedings, volume 1495 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65082-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.Q4F653 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1495.htm;
 http://www.springerlink.
com/content/978-3-540-
65082-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1495.

Atzeni:2004:CMI

[ACL^{+}04] Paolo Atzeni, Wesley Chu, Hongjun Lu, Shuigeng Zhou, and Tok Wang Ling, editors. Conceptual modeling ER 2004: 23rd international conference on conceptual modeling, Shanghai, China, November 8–12, 2004. Proceedings, volume 3288 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23723-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        23723-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3288.

Agrawal:2012:TAM

[ACL12] Manindra Agrawal, S. Barry Cooper, and Angsheng Li, editors. Theory and Applications of Models of Computation: 9th Annual Conference, TAMC 2012, Beijing, China, May 16–21, 2012. Proceedings, volume 7287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29951-2 (print), 3-642-29952-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29952-0.

Aberer:2007:SWI

[ACN^{+}07] Karl Aberer, Key-Sun Choi, Natasha Noy, Dean Allemang, Kyung-Il Lee, Lyndon Nixon, Jennifer Golbeck, Peter Mika, Diana Maynard, Riichiro Mizoguchi, Guus Schreiber, and Philippe Cudré-Mauroux, editors. The Semantic Web: 6th International Semantic Web Conference, 2nd Asian Semantic Web Conference, ISWC 2007 + ASWC 2007, Busan, Korea, November 11–15, 2007. Proceedings, volume 4825 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76297-3 (print), 3-540-76298-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76298-0.

Agosti:2001:LIR

[ACP01] Maristella Agosti, Fabio Crestani, and Gabriella Pasi, editors. Lectures on information retrieval: Third European Summer-School, ESSIR 2000 Varenna, Italy, September 11– 15, 2000, revised lectures, volume 1980 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41933-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1980.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1980.
htm;   http://www.springerlink.
com/content/978-3-540-
41933-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1980.

Apostolico:2005:CPM

[ACP05] Alberto Apostolico, Maxime Crochemore, and Kunsoo Park, editors. Combinatorial pattern matching: 16th annual symposium, CPM 2005, Jeju Island, Korea, June 19– 22, 2005, proceedings, volume 3537 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26201-6 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26201-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3537;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b137128.

Amati:2007:AIR

[ACR07]  Giambattista  Amati,

Claudio Carpineto, and Giovanni Romano, editors. Advances in Information Retrieval: 29th European Conference on IR Research, ECIR 2007, Rome, Italy, April 2– 5, 2007. Proceedings, volume 4425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71494-4 (print), 3-540-71496-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71496-5.

Autexier:2008:ICM

[ACR^{+}08] Serge Autexier, John Campbell, Julio Rubio, Volker Sorge, Masakazu Suzuki, and Freek Wiedijk, editors. Intelligent Computer Mathematics: 9th International Conference, AISC 2008, 15th Symposium, Calculemus 2008, 7th International Conference, MKM 2008, Birmingham, UK, July 28 — August 1, 2008. Proceedings, volume 5144 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85109-7 (print), 3-540-85110-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85110-3.

Atzeni:2012:CMI

[ACR12] Paolo Atzeni, David Cheung, and Sudha Ram, editors. Conceptual Modeling: 31st International Conference ER 2012, Florence, Italy, October 15–18, 2012. Proceedings, volume 7532 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34001-6 (print), 3-642-34002-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
34002-4.

Aarts:2003:AIF

[ACvLdR03] Emile H. L. Aarts, René Collier, Evert van Loenen, and Boris de Ruyter, editors. Ambient Intelligence: First European symposium, EUSAI 2003, Veldhoven, The Netherlands, November 3–4, 2003: Proceedings, volume 2875 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20418-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5915. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2875.htm;
http://www.springerlink.
com/content/978-3-540-
20418-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2875;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94080.

Almasi:2007:LCP

[ACW07] George Almási, Călin Caşcaval, and Peng Wu, editors. Languages and Compilers for Parallel Computing: 19th International Workshop, LCPC 2006, New Orleans, LA, USA, November 2–4, 2006. Revised Papers, volume 4382 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72520-2 (print), 3-540-72521-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72521-3.

Alexandrov:1998:RAP

[AD98] Vassil Alexandrov and J. J. Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 5th European PVM/MPI User’s Group Meeting, Liverpool, UK, September 7–9, 1998: Proceedings, volume 1497 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65041-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1497. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1497.htm;
 http://www.springerlink.
com/content/978-3-540-

65041-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1497.

Jointly sponsored by the Computer Science Dept., University of Liverpool and Oak Ridge National Laboratory.

Ayache:2003:SSS

[AD03] Nicholas Ayache and Herve Delingette, editors. Surgery simulation and soft tissue modeling, volume 2673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40439-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN RD29.7 .I55 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2673.
htm;   http://www.springerlink.
com/content/978-3-540-
40439-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2673.

This book contains the written contributions to the International Symposium on Surgery Simulation and Soft Tissue Modeling (IS4TM 2003).

Abadi:2005:CCT

[AD05] Martín Abadi and Luca De Alfaro, editors. CON-CUR 2005 — Concurrency Theory: 16th International Conference, CONCUR 2005, San Francisco, CA, USA, August 23–26, 2005: Proceedings, volume 3653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28309-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3653;
 http://www.springerlink.
com/content/978-3-540-
28309-6.

Ali:2006:AAA

[AD06] Moonis Ali and Richard Dapoigny, editors. Advances in Applied Artificial Intelligence: 19th International Conference on Industrial, Engineering and Other Applications of Applied Intelligent Systems, IEA/AIE 2006, Annecy, France, June 27–30, 2006. Proceedings, volume 4031 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35453-0 (print), 3-540-35454-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35454-3.

Atluri:2011:CSE

[AD11] Vijay Atluri and Claudia Diaz, editors. Computer Security — ES-ORICS 2011: 16th European Symposium on Research in Computer Security, Leuven, Belgium, September 12–14,2011. Proceedings, volume 6879 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23821-1 (print), 3-642-23822-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23822-2.

Adam:1996:DLR

[Ada96] Nabil R. Adam, editor. Digital libraries: research and technology advances: ADL ’95 Forum, McLean, Virginia, USA, May 15– 17, 1995: Selected Papers, volume 1082 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61410-9 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z699 .D532 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1082.
htm;   http://www.springerlink.
com/content/978-3-540-
61410-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1082.

Althoff:2005:PKM

[ADB^{+}05]  Klaus-Dieter  Althoff,

Andreas Dengel, Ralph Bergmann, Markus Nick, and Thomas Roth-Berghofer, editors. Professional knowledge management: third biennial conference, WM 2005, Kaiserslautern, Germany, April 10–13, 2005: revised selected papers, volume 3782 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30465-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HD30.2 .W6 2005. URL

http://www.springerlink.
com/content/978-3-540-
30465-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3782.

Alexander:2012:EPPa

[ADB^{+}12a] Michael Alexander, Pasqua D’Ambra, Adam Belloum, George Bosilca, Mario Cannataro, Marco Danelutto, Beniamino Di Martino, Michael Gerndt, Emmanuel Jeannot, Raymond Namyst, Jean Roman, Stephen L. Scott, Jesper Larsson Traff, Geoffroy Vallée, and Josef Weidendorfer, editors. Euro-Par 2011: Parallel Processing Workshops: CCPI, CGWS, HeteroPar, HiBB, HPCVirt, HPPC, HPSS, MDGS, ProPer, Resilience, UCHPC, VHPC, Bordeaux, France, August 29 — September 2, 2011, Revised Selected Papers, Part I, volume 7155 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29736-6 (print), 3-642-29737-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29737-3.

Alexander:2012:EPPb

[ADB^{+}12b] Michael Alexander, Pasqua D’Ambra, Adam Belloum, George Bosilca, Mario Cannataro, Marco Danelutto, Beniamino Di Martino, Michael Gerndt, Emmanuel Jeannot, Raymond Namyst, Jean Roman, Stephen L. Scott, Jesper Larsson Traff, Geoffroy Vallée, and Josef Weidendorfer, editors. Euro-Par 2011: Parallel Processing Workshops: CCPI, CGWS, HeteroPar, HiBB, HPCVirt, HPPC, HPSS, MDGS, ProPer, Resilience, UCHPC, VHPC, Bordeaux, France, August 29 — September 2, 2011, Revised Selected Papers, Part II, volume 7156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29739-0 (print), 3-642-29740-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29740-3.

Aichernig:2012:FMC

[AdBB12] Bernhard K. Aichernig, Frank S. de Boer, and Marcello M. Bonsangue, editors. Formal Methods for Components and Objects: 9th International Symposium, FMCO 2010, Graz, Austria, November 29 — December 1, 2010. Revised Papers, volume 6957 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25270-2 (print), 3-642-25271-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25271-6.

Ausiello:1989:ALP

[ADCR89] G. Ausiello, M. Dezani-Ciancaglini, and S. Ronchi Della Rocca, editors. Automata, languages, and programming: 16th international colloquim, Stresa, Italy, July 11– 15, 1989: proceedings, volume 372 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51371-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.372. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0372.htm;
 http://www.springerlink.
com/content/978-0-387-

                                        51371-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=372.

“ICALP 89”–spine.

Agrawal:2008:TAM

[ADDL08] Manindra Agrawal, Dingzhu Du, Zhenhua Duan, and Angsheng Li, editors. Theory and Applications of Models of Computation: 5th International Conference, TAMC 2008, Xi’an, China, April 25–29, 2008. Proceedings, volume 4978 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79227-9 (print), 3-540-79228-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79228-4.

Aceto:2008:ALPa

[ADG^{+}08a] Luca Aceto, Ivan Damgård, Leslie Ann Goldberg, Magnús M. Halldórsson, Anna Ingólfsdóttir, and Igor Walukiewicz, editors. Automata, languages and programming: 35th international colloquium, ICALP 2008, Reykjavik, Iceland, July 7–11, 2008, proceedings, part I, volume 5125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. ISBN 3-540-70574-0 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/7818221152674403/
;   http://www.springerlink.
com/content/978-3-540-
70574-1.

Aceto:2008:ALPb

[ADG^{+}08b] Luca Aceto, Ivan Damgård, Leslie Ann Goldberg, Magnús M. Halldórsson, Anna Ingólfsdóttir, and Igor Walukiewicz, editors. Automata, Languages and Programming: 35th International Colloquium, ICALP 2008, Reykjavik, Iceland, July 7–11, 2008, Proceedings, Part II, volume 5126 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70582-1 (print), 3-540-70583-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70583-3.

Ayache:2012:MICa

[ADGM12a] Nicholas Ayache, Hervé Delingette, Polina Golland, and Kensaku Mori, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2012: 15th International Conference, Nice, France, October 1–5, 2012, Proceedings, Part I, volume 7510 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33414-8 (print), 3-642-33415-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33415-3.

Ayache:2012:MICb

[ADGM12b] Nicholas Ayache, Hervé Delingette, Polina Golland, and Kensaku Mori, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2012: 15th International Conference, Nice, France, October 1–5, 2012, Proceedings, Part II, volume 7511 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33417-2 (print), 3-642-33418-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33418-4.

Ayache:2012:MICc

[ADGM12c] Nicholas Ayache, Hervé Delingette, Polina Golland, and Kensaku Mori, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2012: 15th International Conference, Nice, France, October 1–5, 2012, Proceedings, Part III, volume 7512 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33453-9 (print), 3-642-33454-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33454-2.

Alexandrov:2001:CSIa

[ADJ^{+}01a] Vassil Alexandrov, Jack J.

Dongarra, Benjoe A. Juliano, René S. Renner, and C. J. Kenneth Tan, editors. Computational science — ICCS 2001: International Conference, San Francisco, CA, USA, May 28– 30, 2001: proceedings, Part I, volume 2073 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42232-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2001 v.1-2 (2001); QA267.A1 L43 no.2073-2074 Library has v.1-2. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2073.htm;
 http://link.springer-
ny.com/link/service/
series/0558/tocs/t2074.
htm;   http://www.springerlink.
com/content/978-3-540-
42232-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2073.

Alexandrov:2001:CSIb

[ADJ^{+}01b] Vassil Alexandrov, Jack J. Dongarra, Benjoe A. Juliano, René S. Renner, and C. J. Kenneth Tan, editors. Computational science — ICCS 2001: International Conference, San Francisco, CA, USA, May 28– 30, 2001: proceedings, Part II, volume 2074 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42233-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2001 v.1-2 (2001); QA267.A1 L43 no.2073-2074 Library has v.1-2. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2073.htm;
 http://link.springer-
ny.com/link/service/
series/0558/tocs/t2074.
htm;   http://www.springerlink.
com/content/978-3-540-
42233-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2074.

Aykanat:2004:CIS

[ADK04] Cevdet Aykanat, Tugrul Dayar, and Ibrahim Korpeoğlu, editors. Computer and Information Sciences—ISCIS 2004: 19th International Symposium, Kemer-Antalya, Turkey, October 27– 29, 2004: Proceedings, volume 3280 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23526-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3280.htm;
 http://www.springerlink.
com/content/978-3-540-
23526-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=3280;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b101749.

Andres:2005:DGC

[ADL05] Eric Andres, Guillaume Damiand, and Pascal Lienhardt, editors. Discrete geometry for computer imagery: 12th international conference, DGCI 2005, Poitiers, France, April 13–15, 2005: Proceedings, volume 3429 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25513-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25513-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3429;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b135490.

Avrachenkov:2009:AMW

[ADL09] Konstantin E. Avrachenkov, Debora Donato, and Nelly Litvak, editors. Algorithms and models for the web-graph: 6th international workshop, WAW 2009 Barcelona, Spain, February 12– 13, 2009 proceedings, volume 5427 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-540-95995-5, 3-540-95994-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W39 2009.

URL http://www.springerlink.
com/content/978-3-540-
95995-3.

Andre:2006:PIT

[ADM^{+}06] Elisabeth André, Laila Dybkjær, Wolfgang Minker, Heiko Neumann, and Michael Weber, editors. Perception and Interactive Technologies: International Tutorial and Research Workshop, PIT 2006 Kloster Irsee, Germany, June 19–21, 2006. Proceedings, volume 4021 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34743-7 (print), 3-540-34744-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34744-6.

Andre:2008:PMD

[ADM^{+}08] Elisabeth André, Laila Dybkjær, Wolfgang Minker, Heiko Neumann, Roberto Pieraccini, and Michael Weber, editors. Perception in Multimodal Dialogue Systems: 4th IEEE Tutorial and Research Workshop on Perception and Interactive Technologies for Speech-Based Systems, PIT 2008, Kloster Irsee, Germany, June 16–18, 2008. Proceedings, volume 5078 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69368-8 (print), 3-540-69369-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69369-7.

Agha:2011:FMA

[ADM11] Gul Agha, Olivier Danvy, and José Meseguer, editors. Formal Modeling: Actors, Open Systems, Biological Systems: Essays Dedicated to Carolyn Talcott on the Occasion of Her 70th Birthday, volume 7000 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24932-9 (print), 3-642-24933-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24933-4.

Ardagna:2012:BSM

[ADM^{+}12] Claudio A. Ardagna, Ernesto Damiani, Leszek A. Maciaszek, Michele Missikoff, and Michael Parkin, editors. Business System Management and Engineering: From Open Issues to Applications, volume 7350 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32438-X (print), 3-642-32439-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32439-0.

Andre:2004:ADS

[ADMH04] Elisabeth André, Laila Dybkjær, Wolfgang Minker, and Paul Heisterkamp, editors. Affective Dialogue Systems: Tutorial and Research Workshop, ADS 2004, Kloster Irsee, Germany, June 14–16, 2004: Proceedings, volume 3068 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22143-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 A39 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3068.
htm;   http://www.springerlink.
com/content/978-3-540-
22143-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3068;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98229.

Aldewereld:2009:ESA

[ADP09] Huib Aldewereld, Virginia Dignum, and Gauthier Picard, editors. Engineering Societies in the Agents World X: 10th International Workshop, ESAW 2009, Utrecht, The Netherlands, November 18– 20, 2009. Proceedings, volume 5881 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10202-6 (print), 3-642-10203-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10203-5.

Auer:2011:WEI

[ADP11] Sören Auer, Oscar Díaz, and George A. Papadopoulos, editors. Web Engineering: 11th International Conference, ICWE 2011, Paphos, Cyprus, June 20–24, 2011, volume 6757 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22232-3 (print), 3-642-22233-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22233-7.

Agha:2001:COO

[ADR01] Gul A. Agha, Fiorella De Cindio, and Grzegorz Rozenberg. Concurrent object-oriented programming and Petri Nets: advances in Petri Nets, volume 2001 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41942-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 537 pp. LCCN QA76.64 .C6353 2001; QA267.A1 L43 no.2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2001.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41942-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2001.

Alonso:2007:BPM

[ADR07] Gustavo Alonso, Peter Dadam, and Michael Rosemann, editors. Business Process Management: 5th International Conference, BPM 2007, Brisbane, Australia, September 24– 28, 2007. Proceedings, volume 4714 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75182-3 (print), 3-540-75183-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75183-0.

Avron:2008:PCS

[ADR08] Arnon Avron, Nachum Dershowitz, and Alexander Rabinovich, editors. Pillars of Computer Science: Essays Dedicated to Boris (Boaz) Trakhtenbrot on the Occasion of His 85th Birthday, volume 4800 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78126-9 (print), 3-540-78127-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78127-1.

Ayache:2009:FIM

[ADS09] Nicholas Ayache, Hervé Delingette, and Maxime Sermesant, editors. Functional Imaging and Modeling of the Heart: 5th International Conference, FIMH 2009, Nice, France, June 3–5, 2009. Proceedings, volume 5528 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01931-5 (print), 3-642-01932-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01932-6.

Andersen:2005:DES

[ADW05] Kim Viborg Andersen, John K. Debenham, and Roland R. Wagner, editors. Database and expert systems applications: 16th international conference, DEXA 2005, Copenhagen, Denmark, August 22–26, 2005: proceedings, volume 3588 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28566-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I15592 2005.

                                        URL http://springerlink.
                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3588;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        28566-3.

Ali:2005:IAA

[AE05] Moonis Ali and Floriana Esposito, editors. Innovations in Applied Artificial Intelligence: 18th International Conference on Industrial and Engineering Applications of Artificial Intelligence and Expert Systems, IEA/AIE 2005, Bari, Italy, June 22–24, 2005. Proceedings, volume 3533 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26551-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334. URL http:

//www.springerlink.
com/content/978-3-540-
26551-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3533.

Azar:2006:AEA

[AE06] Yossi Azar and Thomas Erlebach, editors. Algorithms — ESA 2006: 14th Annual European Symposium, Zurich, Switzerland, September 11– 13, 2006. Proceedings, volume 4168 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38875-3 (print), 3-540-38876-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38876-0.

Ahronovitz:1997:DGC

[AF97] Ehoud Ahronovitz and Christophe Fiorio, editors. Discrete geometry for computer imagery: 7th International Workshop, DGCI 97, Montpellier, France, December 1997: proceedings, volume 1347 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63884-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385.D525 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1347.
htm;   http://www.springerlink.
com/content/978-3-540-
63884-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1347.

Alves-Foss:1999:FSS

[AF99a] Jim Alves-Foss, editor. Formal syntax and semantics of Java, volume 1523 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66158-1. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 404 pp. LCCN QA76.73.J38F67 1999. US$49.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1523.htm;
 http://www.springerlink.
com/content/978-3-540-

66158-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1523.

Arikawa:1999:DSS

[AF99b] S. Arikawa and Koichi Furukawa, editors. Discovery science: Second International Conference, DS’99, Tokyo, Japan, December 6– 8, 1999: Proceedings, volume 1721 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66713-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1721. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1721.htm;
http://www.springerlink.
com/content/978-3-540-
66713-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1721.

Alt:2002:SAS

[AF02] Helmut Alt and Afonso Ferreira, editors. STACS 2002: 19th Annual Symposium on Theoretical Aspects of Computer Science, Antibes-Juan les Pins, France, March 14– 16, 2002 proceedings, volume 2285 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43283-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S958 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2285.htm;
 http://www.springerlink.
com/content/978-3-540-
43283-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2285.

Arge:2006:ATS

[AF06] Lars Arge and Rusins Freivalds, editors. Algorithm Theory — SWAT 2006: 10th Scandinavian Workshop on Algorithm Theory, Riga, Latvia, July 6–8, 2006. Proceedings, volume 4059 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35753-X (print), 3-540-35755-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35755-1.

Allbeck:2011:MGI

[AF11] Jan M. Allbeck and Petros Faloutsos, editors. Motion in Games: 4th International Conference, MIG 2011, Edinburgh, UK, November 13–15, 2011. Proceedings, volume 7060 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25089-0 (print), 3-642-25090-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25090-3.

Anacleto:2011:ECI

[AFG^{+}11] Junia Coutinho Anacleto, Sidney Fels, Nicholas Graham, Bill Kapralos, Magy Saif El-Nasr, and Kevin Stanley, editors. Entertainment Computing — ICEC 2011: 10th International Conference, ICEC 2011, Vancouver, Canada, October 5–8, 2011. Proceedings, volume 6972 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24499-8 (print), 3-642-24500-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24500-8.

Andersen:2010:EGI

[AFGvE10] Kim Normann Andersen, Enrico Francesconi, Åke Grönlund, and Tom M. van Engers, editors. Electronic Government and the Information Systems Perspective: First International Conference, EGOVIS 2010, Bilbao, Spain, August 31 — September 2, 2010. Proceedings, volume 6267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15171-X (print), 3-642-15172-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15172-9.

Andersen:2011:EGI

[AFGvE11] Kim Normann Andersen, Enrico Francesconi, Åke Grönlund, and Tom M. van Engers, editors. Electronic Government and the Information Systems Perspective: Second International Conference, EGOVIS 2011, Toulouse, France, August 29 — September 2, 2011. Proceedings, volume 6866 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22960-3 (print), 3-642-22961-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22961-9.

Akl:1990:ACI

[AFK90] Selim G. Akl, F. Fiala, and Waldemar W. Koczkodaj, editors. Advances in computing and information — ICCI ’90, volume 468 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 3-540-53504-7 (Berlin), 0-387-53504-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I578 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0468.htm;
 http://www.springerlink.
com/content/978-3-540-
53504-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=468.

“Organized in cooperation with IEEE Hamilton Section and sponsored by the Natural Sciences and Engineering Research Council of Canada (NSERC), Carleton University in Ottawa, and Laurentian University in Sudbury”– T.p. verso.

Alt:2004:NSR

[AFKL04] René Alt, Andreas Frommer, R. Baker Kearfott, and Wolfram Luther, editors. Numerical Software with Result Verification: International Dagstuhl Seminar, Dagstuhl Castle, Germany, January 19–24, 2003. Revised Papers, volume 2991 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21260-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA297 .N867 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2991.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21260-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2991;
                                         http://www.springeronline.
                                        com/3-540-21260-4.

Anderson:2003:CSR

[AFL03] Stuart Anderson, Massimo Felici, and Beverley Littlewood, editors. Computer Safety, Reliability, and Security: 22nd International Conference, SAFECOMP 2003, Edinburgh, UK, September 23–26, 2003: Proceedings, volume 2788 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20126-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 S23 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2788.
htm;   http://www.springerlink.
com/content/978-3-540-
20126-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2788;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12002.

Althoff:2001:ALS

[AFM01]  Klaus-Dieter  Althoff,

Raimund L. Feldmann, and Wolfgang Müller, editors. Advances in learning software organizations: Third International Workshop, LSO 2001, Kaiserslautern, Germany, September 12– 13, 2001: Proceedings, volume 2176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42574-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .L76 2001; QA267.A1 L43 no.2176. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2176.htm;
 http://www.springerlink.
com/content/978-3-540-
42574-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2176.

Agosti:2010:MMI

[AFP^{+}10] Maristella Agosti, Nicola Ferro, Carol Peters, Maarten de Rijke, and Alan Smeaton, editors. Multilingual and Multimodal Information Access Evaluation: International Conference of the Cross-Language Evaluation Forum, CLEF 2010, Padua, Italy, September 20–23, 2010. Proceedings, volume 6360 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15997-4 (print), 3-642-15998-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15998-5.

Apt:2004:RAC

[AFR^{+}04] Krzysztof R. Apt, François Fages, Francesca Rossi, Péter Szeredi, and Josef Váncza, editors. Recent Advances in Constraints: Joint ERCIM/CoLogNET International Workshop on Constraint Solving and Constraint Logic Programming, CSCLP 2003, Budapest, Hungary, June 30–July 2, 2003. Selected Papers, volume 3010 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21834-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3010.htm;
 http://www.springerlink.
com/content/978-3-540-
21834-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3010;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b96986.

Abiteboul:1989:NRC

[AFS89] S. Abiteboul, P. C. Fischer, and H.-J Schek. Nested relations and complex objects in databases, volume 361 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51171-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 323 pp. LCCN QA76.9.D3 N4731 1989. US$26.90. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0361.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51171-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=361.

“Selected papers from the workshop Theory and applications of nested relations and complex objects held in Darmstadt (West Germany) on April 6–8, 1987”–P. [iii].

Anagnostopoulou:2002:MAI

[AFS02] Christina Anagnostopoulou, Miguel Ferrand, and Alan Smaill, editors. Music and Artificial Intelligence: Second International Conference, IC-MAI 2002, Edinburgh, Scotland, UK, September 12–14, 2002. Proceedings, volume 2445 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44145-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ML73.5 .M84 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2445.htm;
 http://www.springerlink.
com/content/978-3-540-
44145-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2445.

Proceedings of the Second International Conference on Music and Artificial Intelligence, IC-MAI 2002, Edinburgh, Scotland, Sept. 12–14, 2002.

Adriaans:2002:GIA

[AFvZ02] Pieter Adriaans, Henning Fernau, and Menno van Zaanen, editors. Grammatical Inference: Algorithms and Applications: 6th International Colloquium: ICGI 2002, Amsterdam, The Netherlands, September 23–25, 2002. Proceedings, volume 2484 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44239-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .G67 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2484.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44239-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2484.

Albayrak:1998:IAT

[AG98] Sahin Albayrak and Francisco J. Garijo, editors. Intelligent agents for telecommunication applications: Second International Workshop, IATA ’98, Paris, France, July 4–7, 1998: Proceedings, volume 1437 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64720-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1437. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1437.htm;
http://www.springerlink.
com/content/978-3-540-
64720-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1437.

Aardal:2001:IPC

[AG01] Karen Aardal and A. M. H.

Gerards, editors. Integer programming and combinatorial optimization: 8th International IPCO Conference, Utrecht, the Netherlands, June 13– 15, 2001: Proceedings, volume 2081 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42225-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2081; T57.74 .C62 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2081.htm;
http://www.springerlink.
com/content/978-3-540-
42225-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2081.

Aldini:2007:FSA

[AG07] Alessandro Aldini and Roberto Gorrieri, editors. Foundations of Security Analysis and Design IV: FOSAD 2006/ 2007 Tutorial Lectures, volume 4677 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74809-1 (print), 3-540-74810-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74810-6.

Abrial:2009:RMS

[AG09] Jean-Raymond Abrial and Uwe Glässer, editors. Rigorous Methods for Software Construction and Analysis: Essays Dedicated to Egon Börger on the Occasion of His 60th Birthday, volume 5115 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-11446-6 (print), 3-642-11447-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11447-2.

Aldini:2011:FSA

[AG11] Alessandro Aldini and Roberto Gorrieri, editors. Foundations of Security Analysis and Design VI: FOSAD Tutorial Lectures, volume 6858 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23081-4 (print), 3-642-23082-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23082-0.

Abdulrazak:2011:TUS

[AGB^{+}11] Bessam Abdulrazak, Sylvain Giroux, Bruno Bouchard, Hélène Pigot, and Mounir Mokhtari, editors. Toward Useful Services for Elderly and People with Disabilities: 9th International Conference on Smart Homes and Health Telematics, ICOST 2011, Montreal, Canada, June 20– 22, 2011. Proceedings, volume 6719 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21534-3 (print), 3-642-21535-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21535-3.

Ayala:2011:ICS

[AGCP^{+}11] José L. Ayala, Braulio García-Cámara, Manuel Prieto, Martino Ruggiero, and Gilles Sicard, editors. Integrated Circuit and System Design. Power and Timing Modeling, Optimization, and Simulation: 21st International Workshop, PATMOS 2011, Madrid, Spain, September 26–29, 2011. Proceedings, volume 6951 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24153-0 (print), 3-642-24154-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24154-3.

Apostolico:2006:RCM

[AGI^{+}06] Alberto Apostolico, Concettina Guerra, Sorin Istrail, Pavel Pevzner, and Michael Waterman, editors. Research in Computational Molecular Biology: 10th Annual International Conference, RECOMB 2006, Venice, Italy, April 2–5, 2006. Proceedings, volume 3909 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33295-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33295-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3909.

Abramsky:2010:ALP

[AGK^{+}10] Samson Abramsky, Cyril Gavoille, Claude Kirchner, Friedhelm Meyer auf der Heide, and Paul G. Spirakis, editors. Automata, languages and programming: 37th International Colloquium, ICALP 2010, Bordeaux, France, July 6–10, 2010, proceedings, volume 6198–6199 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14164-1 (vol. 1: paperback), 3-642-14161-7 (vol. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I55 2010. URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14164-5/.

Atzori:2006:VCP

[AGL06] Luigi Atzori, Daniel D. Giusto, and Riccardo Leonardi, editors. Visual content processing and representation: 9th international workshop, VLBV 2005, Sardinia, Italy, September 15–16, 2005, revised selected papers, volume 3893 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33578-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33578-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3893.

Alhajj:2007:ADM

[AGL^{+}07] Reda Alhajj, Hong Gao, Xue Li, Jianzhong Li, and Osmar R. Zaïane, editors. Advanced Data Mining and Applications: Third International Conference, ADMA 2007 Harbin, China, August 6–8, 2007 Proceedings, volume 4632 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73870-3 (print), 3-540-73871-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73871-8.

Araki:2003:FFM

[AGM03] Keijiro Araki, Stefania Gnesi, and Dino Mandrioli, editors. FME 2003: Formal Methods: International Symposium of Formal Methods Europe Pisa, Italy, September 8– 14, 2003, volume 2805 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40828-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I593 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2805.
htm;   http://www.springerlink.
com/content/978-3-540-
40828-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2805;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13229.

Aldini:2005:FSA

[AGM05] Alessandro Aldini, Roberto Gorrieri, and Fabio Martinelli, editors. Foundations of Security Analysis and Design III: FOSAD 2004/2005 Tutorial Lectures, volume 3655 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28955-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F685 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28955-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3655.

Adve:2008:LCP

[AGP08] Vikram Adve, María Jesús Garzarán, and Paul Petersen, editors. Languages and Compilers for Parallel Computing: 20th International Workshop, LCPC 2007, Urbana, IL, USA, October 11–13, 2007, Revised Selected Papers, volume 5234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85260-3 (print), 3-540-85261-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85261-2.

Ardizzone:1991:TAI

[AGS91] E. Ardizzone, S. Gaglio, and F. Sorbello, editors. Trends in artificial intelligence: 2nd Congress of the Italian Association for Artificial Intelligence, AI*IA, Palermo, Italy, October 29–31, 1991: proceedings, volume 549 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54712-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .C69 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0549.htm;
http://www.springerlink.
com/content/978-0-387-
54712-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=549.

Antoniou:2011:SWRb

[AGS^{+}11a] Grigoris Antoniou, Marko Grobelnik, Elena Simperl, Bijan Parsia, Dimitris Plexousakis, Pieter De Leenheer, and Jeff Pan, editors. The Semanic Web: Research and Applications: 8th Extended Semantic Web Conference, ESWC 2011, Heraklion, Crete, Greece, May 29 — June 2, 2011, Proceedings, Part II, volume 6644 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21063-5 (print), 3-642-21064-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21064-8.

Antoniou:2011:SWRa

[AGS^{+}11b] Grigoris Antoniou, Marko Grobelnik, Elena Simperl, Bijan Parsia, Dimitris Plexousakis, Pieter De Leenheer, and Jeff Pan, editors. The Semantic Web: Research and Applications: 8th Extended Semantic Web Conference, ESWC 2011, Heraklion, Crete, Greece, May 29–June 2, 2011, Proceedings, Part I, volume 6643 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21033-3 (print), 3-642-21034-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21034-1.

Antoniou:1998:LRC

[AGT98] G. (Grigoris) Antoniou, Aditya Ghose, and Miroslaw Truszczynski, editors. Learning and reasoning with complex representations: PRI-CAI’96 Workshops on Reasoning with Incomplete and Changing Information and on Inducing Complex Representations, Cairns, Australia, August 26–30, 1996: Selected Papers, volume 1359 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64413-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.W67 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1359.
htm;   http://www.springerlink.
com/content/978-3-540-
64413-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1359.

Aguilera:2012:DCI

[Agu12] Marcos K. Aguilera, editor. Distributed Computing: 26th International Symposium, DISC 2012, Salvador, Brazil, October 16–18, 2012. Proceedings, volume 7611 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33650-7 (print), 3-642-33651-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33651-5.

Adleman:1994:ANT

[AH94] Leonard M. Adleman and Ming-Deh A. Huang, editors. Algorithmic number theory: first international symposium, ANTS-I, Ithaca, NY, USA, May 6–9, 1994: proceedings, volume 877 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58691-1 (New York), 3-540-58691-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA241 .A43 1994.

DM66.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0877.htm;
 http://www.springerlink.
com/content/978-0-387-
58691-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=877.

Proceedings of the first Algorithmic Number Theory Symposium.

Alur:1996:CAV

[AH96] Rajeev Alur and T. A.

(Thomas A.) Henzinger, editors. Computer aided verification: 8th international conference, CAV ’96, New Brunswick, NJ, USA, July 31–August 3, 1996: proceedings, volume 1102 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61474-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 C38 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1102.
htm;   http://www.springerlink.
com/content/978-3-540-
61474-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1102.

Apostolico:1997:CPM

[AH97a] Alberto Apostolico and Jotun Hein, editors. Combinatorial pattern matching: 8th Annual Symposium, CPM 97, Aarhus, Denmark, June 30–July 2, 1997: proceedings, volume 1264 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63220-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 S88 1997.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1264.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63220-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1264.

Ausnit-Hood:1997:AQS

[AH97b] Christine Ausnit-Hood, editor. Ada 95 quality and style: Guidelines for Professional Programmers, volume 1344 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63823-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 292 pp. LCCN QA76.73.A35A216 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1344.
htm;   http://www.springerlink.
com/content/978-3-540-
63823-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1344.

Andler:1998:ART

[AH98] Sten F. Andler and Jorgen Hansson, editors. Active, Real-Time, and Temporal Database Systems: Second International Workshop, ARTDB-97, Como, Italy, September 8–9, 1997. Proceedings, volume 1553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65649-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54 .I58 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1553.htm;
http://www.springerlink.
com/content/978-3-540-
65649-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1553.

Aagaard:2000:TPH

[AH00] Mark Aagaard and J. Harrison, editors. Theorem proving in higher order logics: 13th International Conference, TPHOLs 2000, Portland, OR, USA, August 14–18, 2000: Proceedings, volume 1869 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67863-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T655 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1869.htm;
 http://www.springerlink.
com/content/978-3-540-
67863-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1869.

Almeroth:2002:MMI

[AH02] Kevin C. Almeroth and Masum Hasan, editors. Management of Multimedia on the Internet: 5th IFIP/IEEE International Conference on Management of Multimedia Networks and Services, MMNS 2002, Santa Barbara, CA, USA, October 6–9, 2002. Proceedings, volume 2496 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44271-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.75 .I58 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2496.htm;
 http://www.springerlink.
com/content/978-3-540-
44271-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2496.

Alt:2003:SAS

[AH03] Helmut Alt and M. Habib, editors. STACS 2003: 20th Annual Symposium on Theoretical Aspects of Computer Science, Berlin, Germany, February 27–March 1, 2003: Proceedings, volume 2607 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00623-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S956 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2607.htm;
 http://www.springerlink.
com/content/978-3-540-
00623-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2607.

Also available via the World Wide Web.

Ata:2007:MNG

[AH07] Shingo Ata and Choong Seon Hong, editors. Managing Next Generation Networks and Services: 10th Asia-Pacific Network Operations and Management Symposium, AP-NOMS 2007, Sapporo, Japan, October 10–12, 2007. Proceedings, volume 4773 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75475-X (print), 3-540-75476-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75476-3.

Atallah:2010:PET

[AH10] Mikhail J. Atallah and Nicholas J. Hopper, editors. Privacy Enhancing Technologies: 10th International Symposium, PETS 2010, Berlin, Germany, July 21–23, 2010. Proceedings, volume 6205 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14526-4 (print), 3-642-14527-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14527-8.

Apolloni:2007:KBIa

[AHJ07a] Bruno Apolloni, Robert J.

Howlett, and Lakhmi Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 11th International Conference, KES 2007, XVII Italian Workshop on Neural Networks, Vietri sul Mare, Italy, September 12–14, 2007. Proceedings, Part I, volume 4692 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74817-2 (print), 3-540-74819-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74819-9.

Apolloni:2007:KBIb

[AHJ07b] Bruno Apolloni, Robert J.

Howlett, and Lakhmi Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 11th International Conference, KES 2007, XVII Italian Workshop on Neural Networks, Vietri sul Mare, Italy, September 12–14, 2007. Proceedings, Part II, volume 4693 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74826-1 (print), 3-540-74827-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74827-4.

Apolloni:2007:KBIc

[AHJ07c] Bruno Apolloni, Robert J. Howlett, and Lakhmi Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 11th International Conference, KES 2007, XVII Italian Workshop on Neural Networks, Vietri sul Mare, Italy, September 12–14, 2007, Proceedings, Part III, volume 4694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74828-8 (print), 3-540-74829-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74829-8.

Anai:2007:ABS

[AHK07] Hirokazu Anai, Katsuhisa Horimoto, and Temur Kutsia, editors. Algebraic Biology: Second International Conference, AB 2007, Castle of Hagenberg, Austria, July 2–4, 2007. Proceedings, volume 4545 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73432-5 (print), 3-540-73433-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73433-8.

Asada:2008:AAI

[AHMT08] Minoru Asada, John C. T. Hallam, Jean-Arcady Meyer, and Jun Tani, editors. From Animals to Animats 10: 10th International Conference on Simulation of Adaptive Behavior, SAB 2008, Osaka, Japan, July 7– 12, 2008. Proceedings, volume 5040 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69133-2 (print), 3-540-69134-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69134-1.

Ahn:2004:LSO

[Ahn04] Sung Joon Ahn, editor. Least Squares Orthogonal Distance Fitting of Curves and Surfaces in Space, volume 3151 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23966-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3151.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23966-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3151;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104017.

Alur:1996:HSI

[AHS96] Rajeev Alur, T. A. (Thomas A.) Henzinger, and Eduardo D. Sontag, editors. Hybrid systems III: verification and control, volume 1066 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61155-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1066.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1066.
htm;   http://www.springerlink.
com/content/978-3-540-
61155-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1066.

Proceedings of the DI-MACS/SYCON Workshop on Verification and Control of Hybrid Systems, organized October 22–25, 1995 at Rutgers University, New Brunswick, New Jersey.

Aceto:2011:ALPa

[AHS11a] Luca Aceto, Monika Henzinger, and Jiří Sgall, editors. Automata, Languages and Programming: 38th International Colloquium, ICALP 2011, Zurich, Switzerland, July 4–8, 2011, Proceedings, Part I, volume 6755 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22005-3 (print), 3-642-22006-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22006-7.

Aceto:2011:ALPb

[AHS11b] Luca Aceto, Monika Henzinger, and Jiří Sgall, editors. Automata, Languages and Programming: 38th International Colloquium, ICALP 2011, Zurich, Switzerland, July 4–8, 2011, Proceedings, Part II, volume 6756 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22011-8 (print), 3-642-22012-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22012-8.

Atzmueller:2011:ASM

[AHSC11] Martin Atzmueller, Andreas Hotho, Markus Strohmaier, and Alvin Chin, editors. Analysis of Social Media and Ubiquitous Data: International Workshops MSM 2010, Toronto, Canada, June 13, 2010, and MUSE 2010, Barcelona, Spain, September 20, 2010, Revised Selected Papers, volume 6904 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23598-0 (print), 3-642-23599-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23599-3.

Arge:2007:AEA

[AHW07] Lars Arge, Michael Hoffmann, and Emo Welzl, editors. Algorithms — ESA 2007: 15th Annual European Symposium, Eilat, Israel, October 8–10, 2007. Proceedings, volume 4698 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75519-5 (print), 3-540-75520-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75520-3.

Abadi:1997:TAC

[AI97] Martin Abadi and T. (Takayasu) Ito, editors. Theoretical aspects of computer software: International Symposium, TACS ’97, Sendai, Japan, September 23–26, 1997: proceedings, volume 1281 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63388-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN A76.751.T33 1997.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1281.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63388-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1281.

Arnborg:1998:ATS

[AI98] Stefan Arnborg and Lars Ivansson, editors. Algorithm theory — SWAT ’98: 6th Scandinavian Workshop on Algorithm Theory, Stockholm, Sweden, July 8–10, 1998: Proceedings, volume 1432 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64682-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1432.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1432.
htm;   http://www.springerlink.
com/content/978-3-540-
64682-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1432.

Aceto:2006:FSS

[AI06] Luca Aceto and Anna Ingólfsdóttir, editors. Foundations of Software Science and Computation Structures: 9th International Conference, FOSSACS 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 25–31, 2006. Proceedings, volume 3921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33045-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33045-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3921.

Anthony:2012:PTA

[AIL12] Patricia Anthony, Mitsuru Ishizuka, and Dickson Lukose, editors. PRI-CAI 2012: Trends in Artificial Intelligence: 12th Pacific Rim International Conference on Artificial Intelligence, Kuching, Malaysia, September 3–7, 2012. Proceedings, volume 7458 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32694-3 (print), 3-642-32695-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32695-0.

Asano:2011:ACI

[AiNOW11] Takao Asano, Shin ichi Nakano, Yoshio Okamoto, and Osamu Watanabe, editors. Algorithms and Computation: 22nd International Symposium, ISAAC 2011, Yokohama, Japan, December 5–8. Proceedings, volume 7074 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25590-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25590-8.

Arikawa:1994:ALT

[AJ94] S. Arikawa and K. P.

Jantke, editors. Algorithmic learning theory: 4th International Workshop on Analogical and Inductive Inference, AII ’94, 5th International Workshop on Algorithmic Learning Theory, ALT ’94, Reinhardsbrunn Castle, Germany, October 10–15, 1994: proceedings, volume 872 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58520-6 (New York), 3-540-58520-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I615 1994. DM104.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0872.htm;
http://www.springerlink.
com/content/978-0-387-
58520-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=872.

Attali:2001:JSC

[AJ01a] Isabelle Attali and Thomas Jensen, editors. Java on smart cards: programming and security: First International Workshop, JavaCard 2000, Cannes, France, September 14, 2000: Revised Papers, volume 2041 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42167-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.J38 J3635 2000; QA267.A1 L43 no.2041.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2041.
htm;   http://www.springerlink.
com/content/978-3-540-
42167-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2041.

Attali:2001:SCP

[AJ01b] Isabelle Attali and Thomas Jensen, editors. Smart card programming and security: International Conference on Research in Smart Cards, E-smart 2001, Cannes, France, September19-21, 2001: Proceedings, volume 2140 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42610-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.S62 I54 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2140.htm;
http://www.springerlink.
com/content/978-3-540-
42610-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2140.

Apel:2011:SCI

[AJ11] Sven Apel and Ethan Jackson, editors. Software Composition: 10th International Conference, SC 2011, Zurich, Switzerland, June 30 — July 1, 2011. Proceedings, volume 6708 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22044-4 (print), 3-642-22045-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22045-6.

Altinsoy:2009:HAI

[AJB09] M. Ercan Altinsoy, Ute Jekosch, and Stephen Brewster, editors. Haptic and Audio Interaction Design: 4th International Conference, HAID 2009 Dresden, Germany, September 10–11, 2009 Proceedings, volume 5763 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04075-6 (print), 3-642-04076-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04076-4.

Albrecht:1987:PAA

[AJM87] A. Albrecht, H. Jung, and Kurt Mehlhorn, editors. Parallel algorithms and architectures: International Workshop, Suhl, GDR, May 25–30, 1987, proceedings, volume 269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18099-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I6181 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0269.htm;
 http://www.springerlink.
com/content/978-0-387-
18099-1;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=269.

Aura:2012:IST

[AJN12] Tuomas Aura, Kimmo Järvinen, and Kaisa Nyberg, editors. Information Security Technology for Applications: 15th Nordic Conference on Secure IT Systems, NordSec 2010, Espoo, Finland, October 27–29, 2010, Revised Selected Papers, volume 7127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27936-8 (print), 3-642-27937-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27937-9.

Abolmaesumi:2012:IPC

[AJNJ12] Purang Abolmaesumi, Leo Joskowicz, Nassir Navab, and Pierre Jannin, editors. Information Processing in Computer-Assisted Interventions: Third International Conference, IP-CAI 2012, Pisa, Italy, June 27, 2012. Proceedings, volume 7330 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30617-9 (print), 3-642-30618-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30618-1.

Adali:2009:ICA

[AJRB09] Tülay Adali, Christian Jutten, João Marcos Travassos Romano, and Allan Kardec Barros, editors. Independent Component Analysis and Signal Separation: 8th International Conference, ICA 2009, Paraty, Brazil, March 15–18, 2009. Proceedings, volume 5441 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00598-5 (print), 3-642-00599-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00599-2.

Arora:2003:ARC

[AJRS03] Sanjeev Arora, Klaus Jansen, José D. P. Rolim, and Amit Sahai, editors. Approximation, Randomization, and Combinatorial Optimization: Algorithms and Techniques: 6th International Workshop on Approximation Algorithms for Combinatorial Optimization Problems, APPROX 2003, and 7th International Workshop on Randomization and Approximation Techniques in Computer Science, RANDOM 2003, Princeton, NJ, USA, August 24–26, 2003: Proceedings, volume 2764 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40770-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I643 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2764.htm;
http://www.springerlink.
com/content/978-3-540-
40770-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2764;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11961.

Arimura:2000:ALT

[AJS00] Hiroki Arimura, Sanjay Jain, and Arun K. Sharma, editors. Algorithmic Learning Theory:

11th International Conference, ALT 2000, Sydney, Australia, December 11–13, 2000. Proceedings, volume 1968 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41237-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1968. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1968.htm;
 http://www.springerlink.
com/content/978-3-540-
41237-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1968.

Abdallah:2005:CSP

[AJS05] Ali E. Abdallah, Cliff B. Jones, and Jeff W. Sanders, editors. Communicating sequential processes. the first 25 years: symposium on the occasion of 25 years of CSP, London, UK, July 7–8, 2004. Revised invited papers, volume 3525 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25813-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25813-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3525;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136154.

Abiteboul:1990:ITI

[AK90] S. Abiteboul and P. C.

Kanellakis, editors. ICDT ’90: Third International Conference on Database Theory, Paris, France, December 12–14, 1990: proceedings, volume 470 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-53507-1 (Berlin), 0-387-53507-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I557 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0470.htm;
 http://www.springerlink.
com/content/978-3-540-
53507-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=470.

Afrati:1997:DTI

[AK97] Foto Afrati and Phokion Kolaitis, editors. Database theory—ICDT ’97: 6th International Conference, Delphi, Greece, January 8–10, 1997: proceedings, volume 1186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62222-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I557 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1186.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62222-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1186.

Asada:1999:RRS

[AK99] Minoru Asada and Hiroaki Kitano, editors. RoboCup-98: Robot Soccer World Cup II, volume 1604 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66320-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1604.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1604.
htm;   http://www.springerlink.
com/content/978-3-540-
66320-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1604.

Akiyama:2003:DCG

[AK03] Jin Akiyama and Mikio Kano, editors. Discrete and Computational Geometry: Japanese Conference, JCDCG 2002, Tokyo, Japan, December 6–9, 2002: Revised Papers, volume 2866 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20776-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA167 .J33 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2866.htm;
 http://www.springerlink.
com/content/978-3-540-
20776-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2866;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11261.

Abdennadher:2007:RST

[AK07] Nabil Abdennadher and Fabrice Kordon, editors. Reliable Software Technologies — Ada Europe 2007: 12th Ada-Europe International Conference on Reliable Software Technologies, Geneva, Switzerland, June 25– 29, 2007. Proceedings, volume 4498 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73229-2 (print), 3-540-73230-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73230-3.

Aagesen:2010:NSA

[AK10] Finn Arve Aagesen and Svein Johan Knapskog, editors. Networked Services and Applications — Engineering, Control and Management: 16th EUNICE/IFIP WG 6.6 Workshop, EUNICE 2010, Trondheim, Norway, June 28–30, 2010. Proceedings, volume 6164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13970-1 (print), 3-642-13971-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13971-0.

Ardissono:2012:AUM

[AK12] Liliana Ardissono and Tsvi Kuflik, editors. Advances in User Modeling: UMAP 2011 Workshops, Girona, Spain, July 11–15, 2011, Revised Selected Papers, volume 7138 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28508-2 (print), 3-642-28509-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28509-7.

Arun-Kumar:2006:FFS

[AKG06] S. Arun-Kumar and Naveen Garg, editors. FSTTCS 2006: Foundations of Software Technology and Theoretical Computer Science: 26th International Conference, Kolkata, India, December 13–15, 2006. Proceedings, volume 4337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49994-6 (print), 3-540-49995-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49995-4.

Alonso:2003:AAM

[AKK03] Eduardo Alonso, Dimitar Kazakov, and Daniel Kudenko, editors. Adaptive agents and multiagent systems, adaptation and multi-agent learning, volume 2636 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40068-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 A312 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2636.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40068-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2636.

Aberer:2004:DIS

[AKK04] Karl Aberer, Manolis Koubarakis, and Vana Kalogeraki, editors. Databases, Information Systems, and Peer-to-Peer Computing: First International Workshop, DBISP2P 2003: Berlin, Germany, September 7– 8, 2003: Revised Papers, volume 2944 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20968-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.525 .D35 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2944.
htm;   http://www.springerlink.
com/content/978-3-540-
20968-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2944;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95270.

Akl:1995:ADS

[Akl95] Selim G. Akl, editor. Algorithms and data structures: 4th International Workshop, WADS ’95, Kingston, Canada, August 16–18, 1995: proceedings, volume 955 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60220-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 W33 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0955.htm;
 http://www.springerlink.
com/content/978-3-540-
60220-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=955.

Antsaklis:1999:HSV

[AKL^{+}99] Panos J. Antsaklis, Wolf Kohn, Michael Lemmon, Anil Nerode, and Shankar Sastry, editors. Hybrid systems V: Proceedings of the Fifth International Hybrid Systems Workshop, held in Notre Dame, Indiana, September 11–13, 1997, volume 1567 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65643-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38 .H953 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1567.htm;

http://www.springerlink.
com/content/978-3-540-
65643-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1567.

Akman:1987:USP

[Akm87] Varol Akman. Unobstructed shortest paths in polyhedral environments, volume 251 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17629-2 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). 103 pp. LCCN TJ211 .A421 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0251.htm;
http://www.springerlink.
com/content/978-0-387-
17629-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=251.

Based on the author’s thesis (Ph. D.), Rensselaer Polytechnic, 1985.

Akman:2001:MUC

[Akm01] Varol Akman, editor.

Modeling and Using Context: Third International and Interdisciplinary Conference, CONTEXT 2001, Dundee, UK, July 27–30, 2001:

Proceedings, volume 2116 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42379-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2116; QA76.9.C65 I537 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2116.htm;
 http://www.springerlink.
com/content/978-3-540-
42379-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2116.

Aleven:2010:ITSa

[AKM10a] Vincent Aleven, Judy Kay, and Jack Mostow, editors. Intelligent Tutoring Systems: 10th International Conference, ITS 2010, Pittsburgh, PA, USA, June 14–18, 2010, Proceedings, Part I, volume 6094 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13387-8 (print), 3-642-13388-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13388-6.

Aleven:2010:ITSb

[AKM10b] Vincent Aleven, Judy Kay, and Jack Mostow, editors. Intelligent Tutoring Systems: 10th International Conference, ITS 2010, Pittsburgh, PA, USA, June 14–18, 2010, Proceedings, Part II, volume 6095 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13436-X (print), 3-642-13437-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13437-1.

Asano:2003:GMC

[AKR03] T. Asano, Reinhard Klette, and Christian Ronse, editors. Geometry, morphology, and computational imaging: 11th International Workshop on Theoretical Foundations of Computer Vision, Dagstuhl Castle, Germany, April 7– 12, 2002: revised paper, volume 2616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00916-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .I557 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2616.htm;
 http://www.springerlink.
com/content/978-3-540-
00916-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2616.

Azzopardi:2009:AIR

[AKR^{+}09] Leif Azzopardi, Gabriella Kazai, Stephen Robertson, Stefan Rüger, Milad Shokouhi, Dawei Song, and Emine Yilmaz, editors. Advances in Information Retrieval Theory: Second International Conference on the Theory of Information Retrieval, ICTIR 2009 Cambridge, UK, September 10–12, 2009 Proceedings, volume 5766 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04416-6 (print), 3-642-04417-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
04417-5.

Aloni:2012:LLM

[AKR^{+}12] Maria Aloni, Vadim Kimmelman, Floris Roelofsen, Galit W. Sassoon, Katrin Schulz, and Matthijs Westera, editors. Logic, Language and Meaning: 18th Amsterdam Colloquium, Amsterdam , The Netherlands, December 19–21, 2011, Revised Selected Papers, volume 7218 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31481-3 (print), 3-642-31482-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31482-7.

Avanzi:2009:SAC

[AKS09] Roberto Maria Avanzi, Liam Keliher, and Francesco Sica, editors. Selected Areas in Cryptography: 15th International Workshop, SAC 2008, Sackville, New Brunswick, Canada, August 14–15, Revised Selected Papers, volume 5381 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04158-2 (print), 3-642-04159-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04159-4.

Avrithis:2006:SMF

[AKSO06] Yannis Avrithis, Yiannis Kompatsiaris, Steffen Staab, and Noel E. O’Connor, editors. Semantic Multimedia: First International Conference on Semantics and Digital Media Technologies, SAMT 2006, Athens, Greece, December 6– 8, 2006. Proceedings, volume 4306 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49335-2 (print), 3-540-49337-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49337-2.

Akiyama:2006:DCG

[AKT06] Jin Akiyama, Mikio Kano, and Xuehou Tan, editors. Discrete and Computational Geometry: Japanese Conference, JCDCG 2004, Tokyo, Japan, October 8–11, 2004, Revised Selected Papers, volume 3742 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30467-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30467-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3742.

Auweter:2012:IKT

[AKTT12] Axel Auweter, Dieter Kranzlmüller, Amirreza Tahamtan, and A. Min Tjoa, editors. ICT as Key Technology against Global Warming: Second International Conference, ICT-GLOW 2012, Vienna, Austria, September 6, 2012. Proceedings, volume 7453 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32605-6 (print), 3-642-32606-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32606-6.

Akiyama:2001:DCG

[AKU01] J. Akiyama, Mikio Kano, and Masatsugu Urabe, editors. Discrete and computational geometry: Japanese Conference, JCDCG 2000, Tokyo, Japan, November 22– 25, 2000: Revised Papers, volume 2098 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42306-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2098; QA167 .J33 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2098.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42306-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2098.

Abe:2001:ALT

[AKZ01] Naoki Abe, Roni Khardon, and Thomas Zeugmann, editors. Algorithmic Learning Theory: 12th International Conference, ALT 2001, Washington, DC, USA, November 25–28, 2001. Proceedings, volume 2225 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42875-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A48 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2225.htm;
 http://www.springerlink.
com/content/978-3-540-
42875-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2225.

Amir:2001:CPM

[AL01] Amihood Amir and Gad M. Landau, editors. Combinatorial pattern matching: 12th annual symposium, CPM 2001, Jerusalem, Israel, July 1–4, 2001: Proceedings, volume 2089 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42271-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 2001.

URL http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t2089.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        42271-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2089.

Alur:2003:EST

[AL03a] Rajeev Alur and Insup Lee, editors. Embedded Software: Third International Conference, EM-SOFT 2003, Philadelphia, PA, USA, October 13–15, 2003: Proceedings, volume 2855 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20223-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2855.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20223-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2855;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13983.

Amadio:2003:CCT

[AL03b] Roberto Amadio and Denis Lugiez, editors. CON-CUR 2003—Concurrency Theory: 14th International Conference, Marseille, France, September 3–5, 2003: Proceedings, volume 2761 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40753-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2761.htm;
 http://www.springerlink.
com/content/978-3-540-
40753-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2761;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11938.

Alferes:2004:LAI

[AL04] José Júlio Alferes and João Leite, editors. Logics in Artificial Intelligence: 9th European Conference, JELIA 2004, Lisbon, Portugal, September 27–30, 2004, Proceedings, volume 3229 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23242-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3229.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23242-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3229;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100483.

Atzeni:2006:PKI

[AL06] Andrea S. Atzeni and Antonio Lioy, editors. Public Key Infrastructure: Third European PKI Workshop: Theory and Practice, EUROPKI 2006, Turin, Italy, June 19–20, 2006. Proceedings, volume 4043 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35151-5 (softcover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        35151-1.

Abdulla:2011:TAC

[AL11a] Parosh Aziz Abdulla and K. Rustan M. Leino, editors. Tools and Algorithms for the Construction and Analysis of Systems: 17th International Conference, TACAS 2011, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2011, Saarbrücken, Germany, March 26–April 3, 2011. Proceedings, volume 6605 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19834-1 (print), 3-642-19835-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19835-9.

Armando:2011:ARS

[AL11b]  Alessandro  Armando

and Gavin Lowe, editors. Automated Reasoning for Security Protocol Analysis and Issues in the Theory of Security: Joint Workshop, ARSPA-WITS 2010, Paphos, Cyprus,March 27– 28, 2010. Revised Selected Papers, volume 6186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16073-5 (print), 3-642-16074-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16074-5.

Armknecht:2012:RCW

[AL12] Frederik Armknecht and Stefan Lucks, editors. Research in cryptology: 4th Western European Workshop, WEWoRC 2011, Weimar, Germany, July 20–22, 2011, Revised selected papers, volume 7242 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34158-6, 3-642-34159-4 (e-book). LCCN QA76.9.A25 W47 2011.

                                        URL http://www.springerlink.
                                        com/content/978-3-642-
                                        34158-8;   http://www.
                                        springerlink.com/content/
                                        978-3-642-34159-5.

Albayrak:1999:IAT

[Alb99] Sahin Albayrak, editor. Intelligent agents for telecommunication applications: Third International Workshop, IATA’99, Stockholm, Sweden, August 9–10, 1999: Proceedings, volume 1699 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66539-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1699; TK5101.A1 I5559 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1699.htm;
http://www.springerlink.
com/content/978-3-540-
66539-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1699.

Alliot:1996:AEE

[All96] J.-M. (Jean-Marc) Alliot, editor. Artificial evolution: European Conference, AE 95, Brest, France, September 4– 6, 1995: Selected Papers, volume 1063 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61108-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .A77 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1063.htm;
http://www.springerlink.
com/content/978-3-540-
61108-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1063.

Aylett:2010:IST

[ALL^{+}10] Ruth Aylett, Mei Yii Lim, Sandy Louchart, Paolo Petta, and Mark Riedl, editors. Interactive Storytelling: Third Joint Conference on Interactive Digital Storytelling, ICIDS 2010, Edinburgh, UK, November 1–3, 2010. Proceedings, volume 6432 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16637-7 (print), 3-642-16638-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16638-9.

Adelsberger:1995:IMC

[ALM95] Heimo H. Adelsberger, Jiri Lazansky, and V. Marik. Information management in computer integrated manufacturing: a comprehensive guide to state-of-the-art CIM solutions, volume 973 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60286-0. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 663 pp. LCCN TS155.63 .I39 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0973.htm;
http://www.springerlink.
com/content/978-3-540-
60286-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=973.

Alonso:2005:MAI

[Alo05] Gustavo Alonso, editor.

Middleware 2005: ACM/ IFIP/USENIX 6th International Middleware Conference, Grenoble, France, November 28– December 2, 2005: proceedings, volume 3790 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30323-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33

2005. URL http://www.
loc.gov/catdir/toc/
fy0604/2005935949.html;
http://www.springerlink.
com/content/978-3-540-
30323-7.

Alpuente:2011:LBP

[Alp11] María Alpuente, editor. Logic-Based Program Synthesis and Transformation: 20th International Symposium, LOP-STR 2010, Hagenberg, Austria, July 23–25, 2010, Revised Selected Papers, volume 6564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20550-X (print), 3-642-20551-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20551-4.

An:2010:AMT

[ALPH10] Aijun An, Pawan Lingras, Sheila Petty, and Runhe Huang, editors. Active Media Technology: 6th International Conference, AMT 2010, Toronto, Canada, August 28–30, 2010. Proceedings, volume 6335 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15469-7 (print), 3-642-15470-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15470-6.

Anta:2011:PDS

[ALR11] Antonio Fernàndez Anta, Giuseppe Lipari, and Matthieu Roy, editors. Principles of Distributed Systems: 15th International Conference, OPODIS 2011, Toulouse, France, December 13–16, 2011. Proceedings, volume 7109 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25872-7 (print), 3-642-25873-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25873-2.

Akoka:2005:PCM

[ALS^{+}05] Jacky Akoka, Stephen W.

Liddle, Il-Yeol Song, Michela Bertolotto, Isabelle Comyn-Wattiau, Samira Si-Said Cherfi, Willem-Jan vanden Heuvel, Bernhard Thalheim, Manuel Kolp, Paolo Bresciani, Juan Trujillo, Christian Kop, and Heinrich C. Mayr, editors. Perspectives in conceptual modeling: ER 2005 workshops AOIS, BP-UML, CoMoGIS, eCOMO, and QoIS, Klagenfurt, Austria, October 24–28, 2005: proceedings, volume 3770 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29395-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26 I567 2005; QA76.9.D26 P477 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        29395-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3770.

Abramowicz:2011:TSB

[ALS^{+}11] Witold Abramowicz, Ignacio M. Llorente, Mike Surridge, Andrea Zisman, and Julien Vayssière, editors. Towards a Service-Based Internet: 4th European Conference, ServiceWave 2011, Poznan, Poland, October 26–28, 2011. Proceedings, volume 6994 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24754-7 (print), 3-642-24755-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24755-2.

Alt:2001:CDM

[Alt01] Helmut Alt. Computational Discrete Mathematics: Advanced Lectures, volume 2122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42775-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 171 pp. LCCN QA39.3 .C66 2001; QA267.A1 L43 no.2122. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2122.htm;
http://www.springerlink.
com/content/978-3-540-
42775-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2122.

Atallah:2011:FAA

[ALZ11] Mikhail Atallah, Xiang-Yang Li, and Binhai Zhu, editors. Frontiers in Algorithmics and Algorithmic Aspects in Information and Management: Joint International Conference, FAW-AAIM 2011, Jinhua, China, May 28–31, 2011. Proceedings, volume 6681 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21203-4 (print), 3-642-21204-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21204-8.

Abramsky:1991:TPIb

[AM91a] Samson Abramsky and Thomas S. E. Maibaum, editors. TAPSOFT ’91: proceedings of the International Joint Conference on Theory and Practice of Software Development, Brighton, UK, April 8–12, 1991: volume 1: Advances in distributed computing (ADC) and Colloquium on Combining Paradigms for Software Development (CCPSD), volume 494 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-53981-6 (New York: paperback), 3-540-53981-6 (Berlin: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I58 1991 v. 1–2 (c1991). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0494.
htm;   http://www.springerlink.
com/content/978-0-387-
53982-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=494.

Contents: v. 2. Advances in distributed computing (ADC) and colloquium on combining paradigms for software development (CCPSD).

Abramsky:1991:TPIa

[AM91b] Samson Abramsky and Thomas S. E. Maibaum, editors. TAPSOFT ’91: proceedings of the International Joint Conference on Theory and Practice of Software Development, Brighton, UK, April 8–12, 1991: volume 1: Colloquium on Trees in Algebra and Programming (CAAP 91), volume 493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-53982-4 (New York), 3-540-53982-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I58 1991 v. 1–2 (c1991). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0493.
htm;   http://www.springerlink.
com/content/978-0-387-
53982-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=493.

Contents: v. 1. Colloquium on Trees in Algebra and Programming (CAAP ’91).

Alblar:1991:AGA

[AM91c] H. Alblar and B. Melichar, editors. Attribute grammars, applications and systems: international summer school SAGA, Prague, Czechoslovakia, June 4–13, 1991: proceedings, volume 545 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54572-7 (Berlin), 0-387-54572-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .A783 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0545.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54572-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=545.

“International Summer School on Attribute Grammars, Applications and Systems”–Foreword.

Aksit:1997:EOO

[AM97] Mehmet Aksit and Satoshi Matsuoka, editors. ECOOP’97 — object-oriented programming: 11th European Conference, Jyväskylä, Finland, June 9–13, 1997: proceedings, volume 1241 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63089-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1241.htm;
 http://www.springerlink.
com/content/978-3-540-
63089-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1241.

Arikawa:1998:DSF

[AM98] Setsuo Arikawa and Hiroshi Motoda, editors. Discovery Science: First International Conference, DS’98, Fukuoka, Japan, December 14– 16, 1998. Proceedings, volume 1532 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65390-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q174.D57 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1532.
htm;   http://www.springerlink.
com/content/978-3-540-
65390-5;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1532.

Arikawa:2000:DST

[AM00] S. Arikawa and Shinichi Morishita, editors. Discovery Science: Third International Conference, DS 2000, Kyoto, Japan, December 4–6, 2000. Proceedings, volume 1967 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41352-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1967. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1967.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41352-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1967.

Arts:2002:IFL

[AM02] Thomas Arts and Markus Mohnen, editors. Implementation of functional languages: 13th International Workshop, IFL 2001, Stockholm, Sweden, September 24– 26, 2001: Selected Papers, volume 2312 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43537-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .I58 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2312.htm;
 http://www.springerlink.
com/content/978-3-540-
43537-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2312.

Aichernig:2003:FMC

[AM03] Bernhard K. Aichernig and Thomas S. E. Maibaum, editors. Formal Methods at the Crossroads: From panacea to Foundational Support: 10th Anniversary Colloquium of UNU/IIST, the International Institute for Software Technology of the United Nations University, Lisbon, Portugal, March 18– 20, 2002: Revised Papers, volume 2757 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20527-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 F67 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2757.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20527-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2757;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11934.

Apostolico:2004:SPI

[AM04] Alberto Apostolico and Massimo Melucci, editors. String processing and information retrieval: 11th international conference, SPIRE 2004, Padova, Italy, October 5–8, 2004: proceedings, volume 3246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23210-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.T48 S682 2004. URL http://

                                        springerlink.metapress.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3246;   http:
                                        //www.springerlink.
                                        com/content/978-3-540-
                                        23210-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=

0302-9743&volume=3246;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100941.

Auer:2005:LTA

[AM05] Peter Auer and Ron Meir, editors. Learning Theory: 18th Annual Conference on Learning Theory, COLT 2005, Bertinoro, Italy, June 27–30, 2005. Proceedings, volume 3559 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26556-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .C66 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26556-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3559.

Altenkirch:2007:TPP

[AM07] Thorsten Altenkirch and Conor McBride, editors. Types for Proofs and Programs: International Workshop, TYPES 2006, Nottingham, UK, April 18–21, 2006, Revised Selected Papers, volume 4502 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74463-0 (print), 3-540-74464-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74464-1.

Ablayev:2010:CST

[AM10] Farid Ablayev and Ernst W. Mayr, editors. Computer Science — Theory and Applications: 5th International Computer Science Symposium in Russia, CSR 2010, Kazan, Russia, June 16–20, 2010. Proceedings, volume 6072 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13181-6 (print), 3-642-13182-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13182-0.

Amadio:2008:FSS

[Ama08a] Roberto Amadio, editor. Foundations of Software Science and Computational Structures: 11th International Conference, FOSSACS 2008, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2008, Budapest, Hungary, March 29 — April 6, 2008. Proceedings, volume 4962 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78497-7 (print), 3-540-78499-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78499-9.

Amaral:2008:LCP

[Ama08b] José Nelson Amaral, editor. Languages and Compilers for Parallel Computing: 21th International Workshop, LCPC 2008, Edmonton, Canada, July 31 — August 2, 2008, Revised Selected Papers, volume 5335 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89739-9 (print), 3-540-89740-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        89740-8.

Ambriola:2001:SPT

[Amb01] Vincenzo Ambriola, editor. Software process technology: 8th European Workshop, EWSPT 2001, Witten, Germany, June 19–21, 2001: Proceedings, volume 2077 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42264-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 E87 2001; QA267.A1 L43 no.2077. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2077.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42264-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2077.

Andreasen:2002:FQA

[AMCL02] Troels Andreasen, Amihai Motro, Henning Christiansen, and Henrik Legind Larsen, editors. Flexible Query Answering Systems: 5th International Conference, FQAS 2002. Copenhagen, Denmark, October 27–29, 2002. Proceedings, volume 2522 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00074-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.Q4 F653 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2522.htm;
 http://www.springerlink.
com/content/978-3-540-
00074-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2522.

America:1991:EEC

[Ame91a] Pierre America, editor.

ECOOP ’91: European Conference on Object-Oriented Programming, Geneva, Switzerland, July 15–19, 1991: proceedings, volume 512 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54262-0 (Berlin), 0-387-54262-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1991. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0512.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54262-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=512.

America:1991:PDS

[Ame91b] Pierre America, editor. Parallel database systems: PRISMA Workshop, Noordwijk, the Netherlands, September 24–26, 1990: proceedings, volume 503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54132-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P74 1990. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0503.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        54132-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=503.

Alvarez:2012:PPR

[AMGJ12] Luis Alvarez, Marta Mejail, Luis Gomez, and Julio Jacobo, editors. Progress in Pattern Recognition, Image Analysis, Computer Vision, and Applications: 17th Iberoamerican Congress, CIARP 2012, Buenos Aires, Argentina, September 3– 6, 2012. Proceedings, volume 7441 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33274-9 (print), 3-642-33275-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33275-3.

Amin:1998:APR

[Ami98] Adnan Amin, editor. Advances in pattern recognition: joint IAPR International Workshops SSPR’98 and SPR’98, Sydney, Australia, August 11–13, 1998: Proceedings, volume 1451 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64858-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 I59 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1451.htm;
http://www.springerlink.

com/content/978-3-540-
64858-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1451.

Atzeni:1999:WWW

[AMM99] Paolo Atzeni, Giansalvatore Mecca, and Alberto O. Mendelzon, editors. The World Wide Web and databases: International Workshop WebDB’98: Valencia, Spain, March 27–28, 1998: Selected Papers, volume 1590 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65890-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 W43 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1590.
htm;   http://www.springerlink.
com/content/978-3-540-
65890-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1590.

Based on the contributions to the International Workshop on the Web and Databases (WebDB’98), held in Valencia, Spain, March 27 and 28, 1998, in conjunction with the Sixth International Conference on Extending Database Technology (EDBT’98).

Apolloni:2006:NNI

[AMNT06] Bruno Apolloni, Maria Marinaro, Giuseppe Nicosia, and Roberto Tagliaferri, editors. Neural Nets: 16th Italian Workshop on Neural Nets, WIRN 2005, and International Workshop on Natural and Artificial Immune Systems, NAIS 2005, Vietri sul Mare, Italy, June 8–11, 2005, Revised Selected Papers, volume 3931 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33183-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33183-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3931.

Araujo:2009:PRI

[AMPT09] Helder Araujo, Ana Maria Mendonça, Armando J. Pinho, and María Inés Torres, editors. Pattern Recognition and Image Analysis: 4th Iberian Conference, IbPRIA 2009 Póvoa de Varzim, Portugal, June 10–12, 2009 Proceedings, volume 5524 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02171-9 (print), 3-642-02172-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02172-5.

An:2008:FIS

[AMRŚ08] Aijun An, Stan Matwin, Zbigniew W. Raś, and Dominik Ślezak, editors. Foundations of Intelligent Systems: 17th International Symposium, ISMIS 2008 Toronto, Canada, May 20–23, 2008 Proceedings, volume 4994 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68122-1 (print), 3-540-68123-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68123-6.

Abrahamsson:2006:EPA

[AMS06]  Pekka  Abrahamsson,

Michele Marchesi, and Giancarlo Succi, editors. Extreme programming and Agile Processes in Software Engineering: 7th International Conference, XP 2006, Oulu, Finland, June 17–22, 2006, Proceedings, volume 4044 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35094-2 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35094-1.

Albers:2009:ALPa

[AMSM^{+}09a] Susanne Albers, Alberto Marchetti-Spaccamela, Yossi Matias, Sotiris Nikoletseas, and Wolfgang Thomas, editors. Automata, languages and programming: 36th international colloquium, ICALP 2009, Rhodes, Greece, July 5–12, 2009; proceedings, part 1, volume 5555 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-642-02926-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I55 2009. URL http:

//www.springerlink.
com/content/978-3-642-
02926-4;   http://www.
springerlink.com/content/
978-3-642-02926-4/.

Albers:2009:ALPb

[AMSM^{+}09b] Susanne Albers, Alberto Marchetti-Spaccamela, Yossi Matias, Sotiris Nikoletseas, and Wolfgang Thomas, editors. Automata, Languages and Programming: 36th International Colloquium, ICALP 2009, Rhodes, Greece, July 5–12, 2009, Proceedings, Part II, volume 5556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02929-9 (print), 3-642-02930-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02930-1.

Amtrup:1999:IST

[Amt99] Jan Willers Amtrup. Incremental Speech Translation, volume 1735 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66753-9. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 200 pp. LCCN QA76.9.N38 A48 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1735.htm;
http://www.springerlink.
com/content/978-3-540-
66753-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1735.

Apolloni:2003:NNI

[AMT03] Bruno Apolloni, Maria Marinaro, and Roberto Tagliaferri, editors. Neural Nets: 14th Italian Workshop on Neural Nets, WIRN VI-ETRI 2003, Vietri sul Mare, Italy, June 4–7, 2003, volume 2859 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20227-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I85 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2859.htm;
http://www.springerlink.
com/content/978-3-540-
20227-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2859;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13826.

Aksit:2003:OCA

[AMU03] Mehmet Aksit, Mira Mezini, and Rainer Unland, editors. Objects, components, architectures, services, and applications for a networked world: International Conference NetObjectDays, NODe 2002, Erfurt, Germany, October 7–10, 2002: Revised Papers, volume 2591 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00737-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .N63 2002; TK5105.888 .N65 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2591.htm;
 http://www.springerlink.
com/content/978-3-540-
00737-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2591.

Also available via the World Wide Web.

Adams:2007:SAC

[AMW07] Carlisle Adams, Ali Miri, and Michael Wiener, editors. Selected areas in cryptography: 14th international workshop, SAC 2007, Ottawa, Canada, August 16–17, 2007; revised selected papers, volume 4876 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-77360-6, 3-540-77359-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 S22 2007eb. URL http:

//www.springerlink.
com/content/978-3-540-
77360-3.

Ahuja:2009:ROL

[AMZ09] Ravindra K. Ahuja, Rolf H. Möhring, and Christos D. Zaroliagis, editors. Robust and Online Large-Scale Optimization: Models and Techniques for Transportation Systems, volume 5868 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05464-1 (print), 3-642-05465-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
05465-5.

Alagar:1995:AMS

[AN95] Vangalur S. Alagar and M. Nivat, editors. Algebraic methodology and software technology: 4th international conference, AMAST ’95, Montréal, Canada, July 3–7, 1995: proceedings, volume 936 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60043-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I5713 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0936.htm;
 http://www.springerlink.
com/content/978-3-540-
60043-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=936.

Adian:1997:LFC

[AN97] S. I. Adian and Anil Nerode, editors. Logical foundations of computer science: 4th International Symposium, LFCS ’97, Yaroslavl, Russia, July 6–12, 1997: proceedings, volume 1234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63045-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .L43 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1234.htm;
http://www.springerlink.
com/content/978-3-540-
63045-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1234.

Augusto:2006:DSH

[AN06] Juan Carlos Augusto and Chris D. Nugent, editors. Designing Smart Homes: The Role of Artificial Intelligence, volume 4008 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35994-X (print), 3-540-35995-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35995-1.

Artemov:2007:LFC

[AN07] Sergei N. Artemov and Anil Nerode, editors. Logical Foundations of Computer Science: International Symposium, LFCS 2007, New York, NY, USA, June 4–7, 2007. Proceedings, volume 4514 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72732-9 (print), 3-540-72734-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72734-7.

Artemov:2009:LFC

[AN09] Sergei Artemov and Anil Nerode, editors. Logical Foundations of Computer Science: International Symposium, LFCS 2009, Deerfield Beach, FL, USA, January 3– 6, 2009. Proceedings, volume 5407 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-92686-0 (print), 3-540-92687-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92687-0.

Anderson:1994:FSE

[And94] Ross Anderson, editor. Fast software encryption: Cambridge Security Workshop, Cambridge, UK, December 9–11, 1993: proceedings, volume 809 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58108-1, 0-387-58108-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C36 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0809.htm;
 http://www.springerlink.
com/content/978-3-540-
58108-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=809.

Andersson:1995:ADC

[And95] S. I. Andersson, editor. Analysis of dynamical and cognitive systems: advanced course, Stockholm, Sweden, August 9–14, 1993: proceedings, volume 888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-58843-4 (Berlin), 0-387-58843-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 A523 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0888.htm;
 http://www.springerlink.
com/content/978-3-540-
58843-6;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=888.

Anderson:1996:IHF

[And96] Ross Anderson, editor. Information hiding: first International Workshop, Cambridge, UK, May 30–June 1, 1996: proceedings, volume 1174 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61996-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25I5414 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1174.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61996-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1174.

Altmann:2008:GEB

[ANF08] Jörn Altmann, Dirk Neumann, and Thomas Fahringer, editors. Grid Economics and Business Models: 5th International Workshop, GECON 2008, Las Palmas de Gran Canaria, Spain, August 26, 2008. Proceedings, volume 5206 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85484-3 (print), 3-540-85485-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85485-2.

Angeline:1997:EPV

[Ang97] Peter J. Angeline, editor.

Evolutionary programming VI: 6th international conference, EP97, Indianapolis, Indiana, USA, April 13–16, 1997: proceedings, volume 1213 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62788-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1213.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1213.
htm;   http://www.springerlink.
com/content/978-3-540-
62788-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1213.

Held jointly with the 1997 IEEE International Conference on Evolutionary Computation.

Altenkirch:1999:TPP

[ANR99]  Thorsten  Altenkirch,

Wolfgang Naraschewski, and Bernhard Reus, editors. Types for proofs and programs: International Workshop, TYPES ’98, Kloster Irsee, Germany, March 27–31, 1998: Selected Papers, volume 1657 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66537-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T96 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1657.
htm;   http://www.springerlink.
com/content/978-3-540-
66537-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1657.

Aizawa:2004:AMIa

[ANS04a] Kiyoharu Aizawa, Yuichi Nakamura, and Shin’ichi Satoh, editors. Advances in Multimedia Information Processing — PCM 2004: 5th Pacific RIM Conference on Multimedia, Tokyo, Japan, November 30–December 3, 2004. Proceedings, Part I, volume 3331 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23974-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I333 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3331.htm;
 http://www.springerlink.
com/content/978-3-540-
23974-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3331;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104114.

Aizawa:2004:AMIb

[ANS04b] Kiyoharu Aizawa, Yuichi Nakamura, and Shin’ichi Satoh, editors. Advances in Multimedia Information Processing — PCM 2004: 5th Pacific RIM Conference on Multimedia, Tokyo, Japan, November 30–December 3, 2004. Proceedings, Part II, volume 3332 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23977-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I333 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3332.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23977-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3332;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104117.

Aizawa:2004:AMIc

[ANS04c] Kiyoharu Aizawa, Yuichi Nakamura, and Shin’ichi Satoh, editors. Advances in Multimedia Information Processing — PCM 2004: 5th Pacific RIM Conference on Multimedia, Tokyo, Japan, November 30–December 3, 2004. Proceedings, Part III, volume 3333 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23985-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I333 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3333.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23985-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3333;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=

volume&id=doi:10.1007/
b104121.

Allen:2009:CSIa

[ANS^{+}09a] Gabrielle Allen, Jarosław Nabrzyski, Edward Seidel, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational science — ICCS 2009. 9th international conference Baton Rouge, LA, USA, May 25–27, 2009. Proceedings, Part I, volume 5544 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009.

Allen:2009:CSIb

[ANS^{+}09b] Gabrielle Allen, Jarosław Nabrzyski, Edward Seidel, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational science — ICCS 2009. 9th international conference Baton Rouge, LA, USA, May 25–27, 2009. Proceedings, Part II, volume 5545 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009.

Antsaklis:1995:HSI

[Ant95] Panos J. Antsaklis, editor. Hybrid systems II, volume 999 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60472-3 (Softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38 .H93 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0999.htm;
 http://www.springerlink.
com/content/978-3-540-
60472-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=999.

Papers from a workshop held at the Mathematical Sciences Institute (MSI) of Cornell University on October 28–30, 1994.

Antsaklis:1997:HSI

[Ant97] Pano J. Antsaklis, editor. Hybrid systems IV: Papers related to the Fourth International Conference on Hybrid Systems, held in Ithaca, N.Y, October 12–14, 1996, volume 1273 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63358-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38.H95 1997.

URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t1273.
htm;   http://www.springerlink.
com/content/978-3-540-
63358-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1273.

Aagaard:2002:FMC

[AO02] Mark D. Aagaard and John W. O’Leary, editors. Formal Methods in Computer-Aided Design: 4th International Conference, FMCAD 2002, Portland, OR, USA, November 6–8, 2002. Proceedings, volume 2517 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00116-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.65 .F53 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2517.htm;
http://www.springerlink.
com/content/978-3-540-
00116-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2517.

Also available via the World Wide Web.

Arbab:2012:FAC

[AÖ12] Farhad Arbab and Peter Csaba Ölveczky, editors. Formal Aspects of Component Software: 8th International Symposium, FACS 2011, Oslo, Norway, September 14–16, 2011, Revised Selected Papers, volume 7253 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35742-3 (print), 3-642-35743-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35743-5.

Ayache:2007:MICa

[AOM07a] Nicholas Ayache, Sébastien Ourselin, and Anthony Maeder, editors. Medical Image Computing and Computer-Assisted Intervention — MIC-CAI 2007: 10th International Conference, Brisbane, Australia, October 29 — November 2, 2007, Proceedings, Part I, volume 4791 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75756-2 (print), 3-540-75757-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75757-3.

Ayache:2007:MICb

[AOM07b] Nicholas Ayache, Sébastien Ourselin, and Anthony Maeder, editors. Medical Image Computing and Computer-Assisted Intervention — MIC-CAI 2007: 10th International Conference, Brisbane, Australia, October 29 — November 2, 2007, Proceedings, Part II, volume 4792 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75758-9 (print), 3-540-75759-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75759-7.

Artikis:2008:ESA

[AOSV08] Alexander Artikis, Gregory M. P. O’Hare, Kostas Stathis, and George Vouros, editors. Engineering Societies in the Agents World VIII: 8th International Workshop, ESAW 2007, Athens, Greece, October 22–24, 2007, Revised Selected Papers, volume 4995 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87653-7 (print), 3-540-87654-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87654-0.

Ausiello:1983:TAP

[AP83] G. Ausiello and M. Protasi, editors. Trees in algebra and programming: 8th colloquium, L’Aquila, March 9–11, 1983: proceedings, volume 159 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-12727-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166.2 .C3 1983. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0159.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        12727-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=159.

Abdulrab:1993:WER

[AP93] Habib Abdulrab and Jean-Pierre Pecuchet, editors. Word equations and related topics: Second International Workshop, IWW-ERT ’91, Rouen, France, October 7–9, 1991: proceedings, volume 677 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56730-5 (Berlin), 0-387-56730-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 I94 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0677.htm;
 http://www.springerlink.
com/content/978-3-540-
56730-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=677.

Alferes:1996:RLP

[AP96] Jose Julio Alferes and Luis Moniz Pereira. Reasoning with logic programming, volume 1111 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61488-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xx + 326 pp. LCCN QA76.63.A33 1996. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1111.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61488-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1111.

Alur:2004:HSC

[AP04a] Rajeev Alur and George J. Pappas, editors. Hybrid Systems: Computation and Control: 7th International Workshop, HSCC 2004, Philadelphia, PA, USA, March 25–27, 2004: Proceedings, volume 2993 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21259-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38 .H73 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2993.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21259-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2993;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96398.

Alur:2004:CAV

[AP04b] Rajeev Alur and Doron A.

Peled, editors. Computer Aided Verification: 16th International Conference, CAV 2004, Boston, MA, USA, July 13–17, 2004: Proceedings, volume 3114 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22342-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3114.htm;
http://www.springerlink.
com/content/978-3-540-
22342-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3114;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98490.

Arvind:2007:FFS

[AP07] V. Arvind and Sanjiva Prasad, editors. FSTTCS 2007: Foundations of Software Technology and Theoretical Computer Science: 27th International Conference, New Delhi, India, December 12–14, 2007. Proceedings, volume 4855 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77049-6 (print), 3-540-77050-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77050-3.

Abdennadher:2009:AGP

[AP09] Nabil Abdennadher and Dana Petcu, editors. Advances in Grid and Pervasive Computing: 4th International Conference, GPC 2009, Geneva, Switzerland, May 4–8, 2009. Proceedings, volume 5529 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01670-7 (print), 3-642-01671-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01671-4.

Amir:2010:CPM

[AP10] Amihood Amir and Laxmi Parida, editors. Combinatorial Pattern Matching: 21st Annual Symposium, CPM 2010, New York, NY, USA, June 21–23, 2010. Proceedings, volume 6129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13508-0 (print), 3-642-13509-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13509-5.

Antunes:2011:PAI

[AP11] Luis Antunes and H. Sofia Pinto, editors. Progress in Artificial Intelligence: 15th Portuguese Conference on Artificial Intelligence, EPIA 2011, Lisbon, Portugal, October 10–13, 2011. Proceedings, volume 7026 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24768-7 (print), 3-642-24769-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24769-9.

Aluru:2007:HPC

[APBP07] Srinivas Aluru, Manish Parashar, Ramamurthy Badrinath, and Viktor K. Prasanna, editors. High Performance Computing — HiPC 2007: 14th International Conference, Goa, India, December 18–21, 2007. Proceedings, volume 4873 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77219-7 (print), 3-540-77220-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77220-0.

Abdalla:2009:ACN

[APFV09] Michel Abdalla, David Pointcheval, Pierre-Alain Fouque, and Damien Vergnaud, editors. Applied Cryptography and Network Security: 7th International Conference, ACNS 2009, Paris-Rocquencourt, France, June 2–5, 2009. Proceedings, volume 5536 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01956-0 (print), 3-642-01957-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01957-9.

Andrews:2011:CSD

[APHA11] Simon Andrews, Simon Polovina, Richard Hill, and Babak Akhgar, editors. Conceptual Structures for Discovering Knowledge: 19th International Conference on Conceptual Structures, ICCS 2011, Derby, UK, July 25–29, 2011. Proceedings, volume 6828 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22687-6 (print), 3-642-22688-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22688-5.

Audebaud:2008:MPC

[APM08] Philippe Audebaud and Christine Paulin-Mohring, editors. Mathematics of Program Construction: 9th International Conference, MPC 2008, Marseille, France, July 15–18, 2008. Proceedings, volume 5133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70593-7 (print), 3-540-70594-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70594-9.

Antunes:2008:MAB

[APN08] Luis Antunes, Mario Paolucci, and Emma Norling, editors. Multi-Agent-Based Simulation VIII: International Workshop, MABS 2007, Honolulu, HI, USA, May 15, 2007, Revised and Invited Papers, volume 5003 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70915-0 (print), 3-540-70916-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70916-9.

Apostolico:1992:CPM

[Apo92] Alberto Apostolico, editor. Combinatorial pattern matching: third annual symposium, Tucson, Arizona, USA, April 29–May 1, 1992: proceedings, volume 644 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-56024-6 (Berlin), 0-387-56024-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0644.htm;
 http://www.springerlink.
com/content/978-3-540-
56024-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=644.

Apostolico:1993:CPM

[Apo93] Alberto Apostolico, editor. Combinatorial pattern matching: 4th annual symposium, CPM 93, Padova, Italy, June 1993: proceedings, volume 684 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56764-X (Berlin), 0-387-56764-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C653 1993. DM58.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0684.
htm;   http://www.springerlink.
com/content/978-3-540-
56764-6;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=684.

Alferes:1996:LAI

[APO96] Jose Julio Alferes, Luis Moniz Pereira, and Ewa Orlowska, editors. Logics in artificial intelligence: European Workshop, JELIA ’96, Evora, Portugal, September 30– October 3, 1996, proceedings, volume 1126 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61630-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334.E985 1996. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1126.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61630-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1126.

Askoxylakis:2012:IST

[APP12] Ioannis Askoxylakis, Henrich C. Pöhls, and Joachim Posegga, editors. Information Security Theory and Practice. Security, Privacy and Trust in Computing Systems and Ambient Intelligent Ecosystems: 6th IFIP WG 11.2 International Workshop, WISTP 2012, Egham, UK, June 20–22, 2012. Proceedings, volume 7322 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30954-2 (print), 3-642-30955-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30955-7.

Alippi:2009:ANNa

[APPE09a] Cesare Alippi, Marios Polycarpou, Christos Panayiotou, and Georgios Ellinas, editors. Artificial Neural Networks — ICANN 2009: 19th International Conference, Limassol, Cyprus, September 14– 17, 2009, Proceedings, Part I, volume 5768 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04273-2 (print), 3-642-04274-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        04274-4.

Alippi:2009:ANNb

[APPE09b] Cesare Alippi, Marios Polycarpou, Christos Panayiotou, and Georgios Ellinas, editors. Artificial Neural Networks — ICANN 2009: 19th International Conference, Limassol, Cyprus, September 14– 17, 2009, Proceedings, Part II, volume 5769 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04276-7 (print), 3-642-04277-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04277-5.

Aberer:2006:WIS

[APR^{+}06] Karl Aberer, Zhiyong Peng, Elke A. Rundensteiner, Yanchun Zhang, and Xuhui Li, editors. Web Information Systems — WISE 2006: 7th International Conference on Web Information Systems Engineering, Wuhan, China, October 23–26, 2006. Proceedings, volume 4255 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48105-2 (print), 3-540-48107-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48107-2.

Antoniou:2006:AAI

[APS06] Grigoris Antoniou, George Potamias, and Costas Spyropoulos, editors. Advances in Artificial Intelligence: 4th Helenic Conference on AI, SETN 2006, Heraklion, Crete, Greece, May 18– 20, 2006. proceedings, volume 3955 of Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34117-X (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34117-8.

Akar:2008:IOM

[APS08] Nail Akar, Michal Pioro, and Charalabos Skianis, editors. IP Operations and Management: 8th IEEE International Workshop, IPOM 2008, Samos Island, Greece, September 22–26, 2008. Proceedings, volume 5275 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87356-2 (print), 3-540-87357-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87357-0.

Alpigini:2002:RSC

[APSZ02] James J. Alpigini, James F. Peters, Andrzej Skowron, and Ning Zhong, editors. Rough Sets and Current Trends in Computing: Third International Conference, RSCTC 2002, Malvern, PA, USA, October 14–16, 2002. Proceedings, volume 2475 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44274-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .R49 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2475.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44274-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2475.

Apt:2000:NTC

[Apt00] Krzysztof R. Apt, editor. New trends in constraints: joint ERCIM/ Compulog Net workshop, Paphos, Cyprus, October 25–27, 1999: Selected Papers, volume 1865 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67885-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .N49 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1865.htm;
http://www.springerlink.
com/content/978-3-540-
67885-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1865.

Artikis:2009:ESA

[APV09] Alexander Artikis, Gauthier Picard, and Laurent Vercouter, editors. Engineering Societies in the Agents World IX: 9th International Workshop, ESAW 2008, Saint-Etienne, France, September 24–26, 2008, Revised Selected Papers, volume 5485 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02561-7 (print), 3-642-02562-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02562-4.

Anderson:2006:PDS

[APW06] James H. Anderson, Giuseppe Prencipe, and Roger Wattenhofer, editors. Principles of Distributed Systems: 9th International Conference, OPODIS 2005, Pisa, Italy, December 12–14, 2005, Revised Selected Papers, volume 3974 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36321-1 (print), 3-540-36322-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36322-4.

Arvind:1998:FST

[AR98] Vikraman Arvind and R. Ramanujam, editors. Foundations of software technology and theoretical computer science:

18th Conference, Chennai, India, December 17–19, 1998: Proceedings, volume 1530 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65384-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1530. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1530.htm;
 http://www.springerlink.
com/content/978-3-540-
65384-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1530;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b71635.

Aggarwal:1999:ACI

[AR99] Alok Aggarwal and C. Pandu Rangan, editors. Algorithms and computation: 10th International Symposium, ISAAC’99, Chennai, India, December 16–18, 1999: Proceedings, volume 1741 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66916-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1741.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1741.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66916-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1741.

Albers:2004:AEA

[AR04] Susanne Albers and Tomasz Radzik, editors. Algorithms—ESA 2004: 12th annual European symposium, Bergen, Norway, September 14– 17, 2004: Proceedings, volume 3221 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23025-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 2004; QA76 A1 L43 3221. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3221.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        23025-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3221;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=

volume&id=doi:10.1007/
b100428.

Altmann:2010:EGC

[AR10] Jörn Altmann and Omer F.

Rana, editors. Economics of Grids, Clouds, Systems, and Services: 7th International Workshop, GECON 2010, Ischia, Italy, August 31, 2010. Proceedings, volume 6296 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15680-0 (print), 3-642-15681-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15681-6.

AlmeidaeCosta:2007:AAL

[ARC^{+}07] Fernando Almeida e Costa, Luis Mateus Rocha, Ernesto Costa, Inman Harvey, and António Coutinho, editors. Advances in Artificial Life: 9th European Conference, ECAL 2007, Lisbon, Portugal, September 10–14, 2007. Proceedings, volume 4648 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74912-8 (print), 3-540-74913-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74913-4.

Arisawa:2002:CMN

[Ari02] Hiroshi Arisawa, editor. Conceptual modeling for new information systems technologies: ER 2001 workshops: HUMACS, DASWIS, ECOMO, and DAMA, Yokohama, November 27–30, 2002: Revised Papers, volume 2465 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44122-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26 I567 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2465.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44122-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2465.

Armando:2002:FCS

[Arm02] Alessandro Armando, editor. Frontiers of combining systems: 4th International Workshop, FroCoS 2002, Santa Margherita Ligure, Italy, April 8–10, 2002: Proceedings, volume 2309 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43381-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.A1 F73 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2309.htm;
 http://www.springerlink.
com/content/978-3-540-
43381-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2309.

Arnold:1990:CCT

[Arn90] A. Arnold, editor. CAAP ’90: 15th Colloquium on Trees in Algebra and Programming, Copenhagen, Denmark, May 1990, proceedings, volume 431 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52590-4 (Berlin), 0-387-52590-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166.2 .C65 1990. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0431.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        52590-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=431.

Abdallah:2003:FAS

[ARS03] Ali E. Abdallah, Peter Ryan, and Steve A. Schneider, editors. Formal Aspects of Security: First International Conference, FASec 2002, London, UK, December 16–18, 2002, Revised Papers, volume 2629 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20693-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F35 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2629.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20693-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2629;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/

b94774. This proceedings volume is dedicated to the memory of Roger Needham.

Abdelzaher:2009:PDS

[ARS09] Tarek Abdelzaher, Michel Raynal, and Nicola Santoro, editors. Principles of Distributed Systems: 13th International Conference, OPODIS 2009, Nîmes, France, December 15–18, 2009. Proceedings, volume 5923 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10876-8 (print), 3-642-10877-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10877-8.

Adams:2009:AID

[ARSB09] Niall M. Adams, Céline Robardet, Arno Siebes, and Jean-François Boulicaut, editors. Advances in Intelligent Data Analysis VIII: 8th International Symposium on Intelligent Data Analysis, IDA 2009, Lyon, France, August 31 — September 2, 2009. Proceedings, volume 5772 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03914-6 (print), 3-642-03915-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03915-7.

Abiteboul:1994:ALP

[AS94] S. Abiteboul and Eli Shamir, editors. Automata, languages and programming: 21st international colloquium, ICALP 94, Jerusalem, Israel, July 11–14, 1994: proceedings, volume 820 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58201-0 (Berlin), 0-387-58201-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.820. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0820.htm;
 http://www.springerlink.
com/content/978-3-540-
58201-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=820.

ICALP (International Colloquium on Automata, Languages, and Programming) is sponsored by the European Association for Theoretical Computer Science.

Antoniou:1998:ATA

[AS98] G. Antoniou and J. K.

Slaney, editors. Advanced topics in artificial intelligence: 11th Australian Joint Conference on Artificial Intelligence, AI’98 Brisbane, Australia, July 13– 17, 1998: Selected Papers, volume 1502 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65138-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1502. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1502.htm;
http://www.springerlink.
com/content/978-3-540-
65138-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1502.

Agouris:1999:ISD

[AS99] Peggy Agouris and Anthony Stefanidis, editors. Integrated spatial databases: digital images and GIS: International Workshop ISD’99, Portland, ME, USA, June 14–16, 1999: Selected Papers, volume 1737 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66931-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55862 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1737.htm;
 http://www.springerlink.
com/content/978-3-540-
66931-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1737.

Agrawal:2002:FTF

[AS02a] Manindra Agrawal and Anil Seth, editors. FST TCS 2002: Foundations of Software Technology and Theoretical Computer Science: 22nd Conference Kanpur, India, December 12–14, 2002. Proceedings, volume 2556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00225-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F68 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2556.htm;

http://www.springerlink.
com/content/978-3-540-
00225-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2556.

Arikawa:2002:PDS

[AS02b] Setsuo Arikawa and Ayumi Shinohara, editors. Progress in discovery science: final report of the Japanese Dicsovery [sic] Science project, volume 2281 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43338-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 684 pp. LCCN QA76.9.D343 P76 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2281.htm;
http://www.springerlink.
com/content/978-3-540-
43338-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2281.

Albrecht:2003:SAF

[AS03] Andreas Albrecht and Kathleen Steinhöfel, editors. Stochastic Algorithms: Foundations and Applications: Second International Symposium, SAGA 2003, Hatfield, UK, September 22–23, 2003: Proceedings, volume 2827 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20103-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2827.htm;
 http://www.springerlink.
com/content/978-3-540-
20103-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2827;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13596.

Alvarez:2006:EAI

[AS06] Carme Alvarez and Maria Serna, editors. Experimental algorithms: 5th international workshop, WEA 2006, Cala Galdana, Menorca, Spain, May 24–27, 2006, proceedings, volume 4007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34597-3 (softcover). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
34597-8.

Arbab:2007:ISF

[AS07a] Farhad Arbab and Marjan Sirjani, editors. International Symposium on Fundamentals of Software Engineering: International Symposium, FSEN 2007, Tehran, Iran, April 17–19, 2007. Proceedings, volume 4767 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75697-3 (print), 3-540-75698-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75698-9.

Arenas:2007:DPL

[AS07b] Marcelo Arenas and Michael I. Schwartzbach, editors. Database Programming Languages: 11th International Symposium, DBPL 2007, Vienna, Austria, September 23–24, 2007, Revised Selected Papers, volume 4797 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75986-7 (print), 3-540-75987-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75987-4.

Azemard:2007:ICS

[AS07c] Nadine Azémard and Lars Svensson, editors. Integrated Circuit and System Design. Power and Timing Modeling, Optimization and Simulation: 17th International Workshop, PAT-MOS 2007, Gothenburg, Sweden, September 3– 5, 2007. Proceedings, volume 4644 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74441-X (print), 3-540-74442-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74442-9.

Arbab:2010:FSE

[AS10] Farhad Arbab and Marjan Sirjani, editors. Fundamentals of Software Engineering: Third IPM International Conference, FSEN 2009, Kish Island, Iran, April 15–17, 2009, Revised Selected Papers, volume 5961 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11622-1 (print), 3-642-11623-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11623-0.

Abreu:2011:ADP

[AS11a] Salvador Abreu and Dietmar Seipel, editors. Applications of Declarative Programming and Knowledge Management: 18th International Conference, INAP 2009, Évora, Portugal, November 3–5, 2009, Revised Selected Papers, volume 6547 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20588-7 (print), 3-642-20589-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20589-7.

Altman:2011:NPC

[AS11b] Erik Altman and Weisong Shi, editors. Network and Parallel Computing: 8th IFIP International Conference, NPC 2011, Changsha, China, October 21–23, 2011. Proceedings, volume 6985 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24402-5 (print), 3-642-24403-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24403-2.

Arbab:2012:FSE

[AS12] Farhad Arbab and Marjan Sirjani, editors. Fundamentals of Software Engineering: 4th IPM International Conference, FSEN 2011, Tehran, Iran, April 20–22, 2011, Revised Selected Papers, volume 7141 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29319-0 (print), 3-642-29320-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
29320-7.

Asano:1990:AP

[Asa90] T. Asano, editor. Algorithms: proceedings, volume 450 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 3-540-52921-7 (Berlin), 0-387-52921-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .I58 1990.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0450.
htm;   http://www.springerlink.
com/content/978-3-540-
52921-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=450.

Asano:1996:ACI

[Asa96] T. Asano, editor. Algorithms and computation: 7th international symposium, ISAAC ’96, Osaka, Japan, December 16–18, 1996: proceedings, volume 1178 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-62048-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I73 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1178.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        62048-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1178.

Asano:1999:CCA

[Asa99] Takao Asano, editor. Computing and combinatorics: 5th annual International Conference, COCOON ’99, Tokyo, Japan, July 26–28, 1999: Proceedings, volume 1627 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66200-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .C62 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1627.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66200-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1627.

Asano:2006:ACI

[Asa06] Tetsuo Asano, editor. Algorithms and Computation: 17th International Symposium, ISAAC 2006, Kolkata, India, December 18–20, 2006. Proceedings, volume 4288 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49694-7 (print), 3-540-49696-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49696-0.

Aspnes:2007:DCS

[ASAM07] James Aspnes, Christian Scheideler, Anish Arora, and Samuel Madden, editors. Distributed Computing in Sensor Systems: Third IEEE International Conference, DCOSS 2007, Santa Fe, NM, USA, June 18– 20, 2007. Proceedings, volume 4549 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73089-3 (print), 3-540-73090-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73090-3.

Akyildiz:2007:NAH

[ASE^{+}07] Ian F. Akyildiz, Raghupathy Sivakumar, Eylem Ekici, Jaudelice Cavalcante de Oliveira, and Janise McNair, editors. NETWORKING 2007. Ad Hoc and Sensor Networks, Wireless Networks, Next Generation Internet: 6th International IFIP-TC6 Networking Conference, Atlanta, GA, USA, May 14–18, 2007. Proceedings, volume 4479 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72605-5 (print), 3-540-72606-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72606-7.

Ambos-Spies:2009:MTC

[ASLM09] Klaus Ambos-Spies, Benedikt Löwe, and Wolfgang Merkle, editors. Mathematical Theory and Computational Practice: 5th Conference on Computability in Europe, CiE 2009, Heidelberg, Germany, July 19–24, 2009. Proceedings, volume 5635 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03072-6 (print), 3-642-03073-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03073-4.

Asonov:2004:QDP

[Aso04] Dmitri Asonov. Querying databases privately: a new approach to private information retrieval, volume 3128 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22441-6. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 113 pp. LCCN QA76.9.A25 A827 2004. URL http:

//www.springerlink.
com/content/978-3-540-
22441-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3128;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/

b98671. Thesis (doctoral), Humboldt Universität, Berlin, 2003.

Alford:1985:DSM

[ASP85] M. W. Alford, H. J.

Siegert, and Manfred Paul. Distributed systems: methods and tools for specification: an advanced course, volume 190 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15216-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 573 pp. LCCN QA76.9.D5 D63 1985. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0190.htm;
 http://www.springerlink.
com/content/978-0-387-
15216-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=190.

Prepared for and presented during the Advanced course on distributed systems, methods and tools for specification, held April 3–12, 1984 and repeated April 16–25, 1985, organized by the Institut für Informatik, Technische Universität München.

Asplund:1998:RST

[Asp98] Lars Asplund, editor. Reliable software technologies — Ada-Europe: 1998 Ada-Europe International Conference on Reliable Software Technologies, Uppsala, Sweden, June 8–12, 1998: Proceedings, volume 1411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64536-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1411. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1411.htm;
 http://www.springerlink.
com/content/978-3-540-
64536-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1411.

Al-Shaer:2001:MMI

[ASP01] Ehab S. Al-Shaer and Giovanni Pacifici, editors. Management of multimedia on the Internet: 4th IFIP/IEEE International Conference on Management of Multimedia Networks and Services, MMNS 2001, Chicago, IL, USA, October 29–November 1, 2001: proceedings, volume 2216 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42786-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2216; TK5103.75 .I58 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2216.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42786-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2216.

An:2007:RSF

[ASR^{+}07] Aijun An, Jerzy Stefanowski, Sheela Ramanna, Cory J. Butz, Witold Pedrycz, and Guoyin Wang, editors. Rough Sets, Fuzzy Sets, Data Mining and Granular Computing: 11th International Conference, RSFDGrC 2007, Toronto, Canada, May 14–16, 2007. Proceedings, volume 4482 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72529-6 (print), 3-540-72530-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72530-5.

Acquisti:2010:TTC

[ASS10] Alessandro Acquisti, Sean W. Smith, and Ahmad-Reza Sadeghi, editors. Trust and Trustworthy Computing: Third International Conference, TRUST 2010, Berlin, Germany, June 21–23, 2010. Proceedings, volume 6101 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13868-3 (print), 3-642-13869-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13869-0.

Adi:2005:RRM

[AST05] Asaf Adi, Suzette Stoutenburg, and Said Tabet, editors. Rules and Rule Markup Languages for the Semantic Web: First International Conference, RULEML 2005, Galway, Ireland, November 10–12, 2005. Proceedings, volume 3791 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29922-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29922-6;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3791.

Agosti:2002:RAT

[AT02a] Maristella Agosti and C. Thanos, editors. Research and Advances Technology for Digital Technology: 6th European Conference, ECDL 2002, Rome, Italy, September 16– 18, 2002. Proceedings, volume 2458 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44178-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z692.C65 E28 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2458.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44178-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2458.

Also available via the

World Wide Web.

Apostolico:2002:CPM

[AT02b] Alberto Apostolico and Masayuki Takeda, editors. Combinatorial Pattern Matching: 13th Annual Symposium, CPM 2002 Fukuoka, Japan, July 3–5, 2002. Proceedings, volume 2373 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43862-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2373.htm;
 http://www.springerlink.
com/content/978-3-540-
43862-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2373.

Arbab:2002:CML

[AT02c] F. (Farhad) Arbab and Carolyn L. Talcott, editors. Coordination models and languages: 5th International Conference, COORDINATION 2002, York, UK, April 8–11, 2002: Proceedings, volume 2315 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43410-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .C695 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2315.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        43410-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2315.

Antunes:2007:MAB

[AT07] Luis Antunes and Keiki Takadama, editors. Multi-Agent-Based Simulation VII: International Workshop, MABS 2006, Hakodate, Japan, May 8, 2006, Revised and Invited Papers, volume 4442 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76536-0 (print), 3-540-76539-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76539-4.

Aoki:2012:FMS

[AT12] Toshiaki Aoki and Kenji Taguchi, editors. Formal Methods and Software Engineering: 14th International Conference on Formal Engineering Methods, ICFEM 2012, Kyoto, Japan, November 12–16, 2012. Proceedings, volume 7635 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34280-9 (print), 3-642-34281-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34281-3.

Aroyo:2009:SWR

[ATC^{+}09] Lora Aroyo, Paolo Traverso, Fabio Ciravegna, Philipp Cimiano, Tom Heath, Eero Hyvönen, Riichiro Mizoguchi, Eyal Oren, Marta Sabou, and Elena Simperl, editors. The Semantic Web: Research and Applications: 6th European Semantic Web Conference, ESWC 2009 Heraklion, Crete, Greece, May 31–June 4, 2009 Proceedings, volume 5554 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02120-4 (print), 3-642-02121-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02121-3.

Atluri:2008:DAS

[Atl08] Vijay Atluri, editor. Data and Applications Security XXII: 22nd Annual IFIP WG 11.3 Working Conference on Data and Applications Security London, UK, July 13–16, 2008 Proceedings, volume 5094 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70566-X (print), 3-540-70567-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70567-3.

Amir:2009:SPI

[ATM09] Amihood Amir, Andrew Turpin, and Alistair Moffat, editors. String Processing and Information Retrieval: 15th International Symposium, SPIRE 2008, Melbourne, Australia, November 10– 12, 2008. Proceedings, volume 5280 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-89096-3 (print), 3-540-89097-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89097-3.

Andrews:2009:AIS

[ATO^{+}09] Paul S. Andrews, Jon Timmis, Nick D. L. Owens, Uwe Aickelin, Emma Hart, Andrew Hone, and Andy M. Tyrrell, editors. Artificial Immune Systems: 8th International Conference, ICARIS 2009, York, UK, August 9– 12, 2009. Proceedings, volume 5666 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03245-1 (print), 3-642-03246-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03246-2.

Atzeni:1993:LDD

[Atz93] Paolo Atzeni. LOGI-DATA+: deductive databases with complex objects, volume 701 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56974-X (Berlin), 0-387-56974-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 273 pp. LCCN QA76.9.D3 L643 1993.

US$44.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0701.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        56974-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=701.

Aucsmith:1998:IHI

[Auc98] David Aucsmith, editor. Information hiding II: Second International Workshop, IH’98: Portland, Oregon, USA, April 1998: Proceedings, volume 1525 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65386-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25I48 1998.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1525.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65386-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1525.

Ausiello:1986:IIC

[Aus86] G. Ausiello, editor. ICDT ’86: International Conference on Database Theory, Rome, Italy, September 8–10, 1986: proceedings, volume 243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-17187-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I558241 1986.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0243.
htm;   http://www.springerlink.
com/content/978-0-387-
17187-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=243.

Aussenac:1993:KAK

[Aus93] N. Aussenac, editor.

Knowledge acquisition for knowledge-based systems: 7th European workshop, EKAW’93, Toulouse and Caylus, France, September 6– 10, 1993: proceedings, volume 723 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-57253-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E36 1993. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0723.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57253-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=723.

Abiteboul:1999:RAT

[AV99] S. (Serge) Abiteboul and Anne-Marie Vercoustre, editors. Research and advanced technology for digital libraries: Third European Conference, ECDL’99, Paris, France, September 22–24, 1999: Proceedings, volume 1696 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66558-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4080 .E28 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1696.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66558-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1696.

Alkassar:2007:VIF

[AV07] Ammar Alkassar and Melanie Volkamer, editors. E-Voting and Identity: First International Conference, VOTE-ID 2007, Bochum, Germany, October 4–5, 2007, Revised Selected Papers, volume 4896 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77492-0 (print), 3-540-77493-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77493-8.

Alpuente:2008:SAI

[AV08] María Alpuente and Germán Vidal, editors. Static Analysis: 15th International Symposium, SAS 2008, Valencia, Spain, July 16–18, 2008. Proceedings, volume 5079 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69163-4 (print), 3-540-69166-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69166-2.

Alexandrov:2006:CSIa

[AvAS06] Vassil N. Alexandrov, Geert Dick van Albada, and Peter M. A. Sloot, editors. Computational Science: ICCS 2006: 6th International Conference, Reading, UK, May 28–31, 2006, Proceedings, Part I, volume 3991 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34379-2 (softcover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34379-0.

Alexandrov:2006:CSIb

[AvASD06a] Vassil N. Alexandrov, Geert Dick van Albada, Peter M. A. Sloot, and Jack Dongarra, editors. Computational Science — ICCS 2006: 6th International Conference, Reading, UK, May 28– 31, 2006. Proceedings, Part II, volume 3992 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34381-4 (print), 3-540-34382-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34382-0.

Alexandrov:2006:CSIc

[AvASD06b] Vassil N. Alexandrov, Geert Dick van Albada, Peter M. A. Sloot, and Jack Dongarra, editors. Computational Science — ICCS 2006: 6th International Conference, Reading, UK, May 28– 31, 2006. Proceedings, Part III, volume 3993 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34383-0 (print), 3-540-34384-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34384-4.

Alexandrov:2006:CSId

[AvASD06c] Vassil N. Alexandrov, Geert Dick van Albada, Peter M. A. Sloot, and Jack Dongarra, editors. Computational Science — ICCS 2006: 6th International Conference, Reading, UK, May 28– 31, 2006, Proceedings, Part IV, volume 3994 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34385-7 (print), 3-540-34386-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        34386-8.

Avgeriou:2012:SER

[Avg12] Paris Avgeriou, editor. Software Engineering for Resilient Systems: 4th International Workshop, SERENE 2012, Pisa, Italy, September 27– 28, 2012. Proceedings, volume 7527 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33175-0 (print), 3-642-33176-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33176-3.

Aarts:1991:PPAa

[AvLR91a] Emile H. L. Aarts, Jan van Leeuwen, and M. Rem, editors. PARLE ’91: parallel architectures and languages Europe: Eindhoven, the Netherlands, June 10–13, 1991: Proceedings. Part I. Parallel Architectures and Algorithms, volume 505 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54151-9 (New York), 3-540-54151-9 (Berlin).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .C61926 1991 v. 1–2 (c1991). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0505.
htm;   http://www.springerlink.
com/content/978-0-387-
54151-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=505.

Aarts:1991:PPAb

[AvLR91b] Emile H. L. Aarts, Jan van Leeuwen, and M. Rem, editors. PARLE ’91: parallel architectures and languages Europe: Eindhoven, the Netherlands, June 10– 13, 1991: Proceedings. Part II. Parallel Languages, volume 506 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54152-7 (New York), 3-540-54152-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .C61926 1991 v. 1–2 (c1991). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0506.
htm;   http://www.springerlink.
com/content/978-0-387-
54152-5;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=506.

Akehurst:2007:MDA

[AVP07] David H. Akehurst, Régis Vogel, and Richard F. Paige, editors. Model Driven Architecture — Foundations and Applications: Third European Conference, ECMDA-FA 2007, Haifa, Israel, June 11–15, 2007, Proceedings, volume 4530 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72900-3 (print), 3-540-72901-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72901-3.

Almeida:1990:NNE

[AW90] L. B. Almeida and C. J. Wellekens, editors. Neural networks: EURASIP Workshop 1990, Sesimbra, Portugal, February 15–17, 1990, proceedings, volume 412 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-52255-7 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .E88 1990. DM42.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0412.
htm;   http://www.springerlink.
com/content/978-0-387-
52255-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=412.

Aha:2001:CBR

[AW01] David W. Aha and Ian D.

Watson, editors. Casebased reasoning research and development: 4th International Conference on Case-Based Reasoning, ICCBR 2001, Vancouver, BC, Canada, July 30–August 2, 2001: Proceedings, volume 2080 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42358-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1; QA76.76.E95 I568 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2080.htm;
 http://www.springerlink.
com/content/978-3-540-
42358-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2080.

Amyot:2005:SAM

[AW05] Daniel Amyot and Alan W. Williams. System analysis and modeling: modeling and analysis of reactive, distributed, and real-time systems, volume 3319 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24561-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24561-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3319.

Agudo:2012:CBR

[AW12] Belén Díaz Agudo and Ian Watson, editors. Case-Based Reasoning Research and Development: 20th International Conference, IC-CBR 2012, Lyon, France, September 3–6, 2012. Proceedings, volume 7466 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32985-3 (print), 3-642-32986-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32986-9.

Aroyo:2011:SWIa

[AWA^{+}11a] Lora Aroyo, Chris Welty, Harith Alani, Jamie Taylor, Abraham Bernstein, Lalana Kagal, Natasha Noy, and Eva Blomqvist, editors. The Semantic Web — ISWC 2011: 10th International Semantic Web Conference, Bonn, Germany, October 23–27, 2011, Proceedings, Part I, volume 7031 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25072-6 (print), 3-642-25073-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25073-6.

Aroyo:2011:SWIb

[AWA^{+}11b] Lora Aroyo, Chris Welty, Harith Alani, Jamie Taylor, Abraham Bernstein, Lalana Kagal, Natasha Noy, and Eva Blomqvist, editors. The Semantic Web — ISWC 2011: 10th International Semantic Web Conference, Bonn, Germany, October 23–27, 2011, Proceedings, Part II, volume 7032 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25092-0 (print), 3-642-25093-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25093-4.

Adam:1996:ECC

[AY96] Nabil R. Adam and Yelena Yesha. Electronic commerce: current research issues and applications, volume 1028 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60738-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 155 pp. LCCN HF5548.33 .A3 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1028.htm;
 http://www.springerlink.
com/content/978-3-540-
60738-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1028.

Ayache:1995:CVV

[Aya95] Nicholas Ayache, editor. Computer vision, virtual reality, and robotics in medicine: first international conference, CVRMed ’95, Nice, France, April 3–6, 1995: proceedings, volume 905 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 0-387-59120-6 (New York), 3-540-59120-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R859.7.C67 I58 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0905.htm;
http://www.springerlink.
com/content/978-0-387-
59120-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=905.

Contributions to the program of the First International Conference on Computer Vision, Virtual Reality, and Robotics in Medicine (CVRMed ’95).

Adam:1995:DLC

[AYB95] Nabil R. Adam, Yelena Yesha, and Bharat K. Bhargava, editors. Digital libraries: current issues: digital libraries workshop DL ’94, Newark NJ, USA, May 19– 20, 1994: Selected Papers, volume 916 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59282-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z699 .D53 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0916.htm;
 http://www.springerlink.
com/content/978-3-540-
59282-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=916.

Andreasen:2009:FQA

[AYB^{+}09] Troels Andreasen, Ronald R. Yager, Henrik Bulskov, Henning Christiansen, and Henrik Legind Larsen, editors. Flexible Query Answering Systems: 8th International Conference, FQAS 2009, Roskilde, Denmark, October 26– 28, 2009. Proceedings, volume 5822 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04956-7 (print), 3-642-04957-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04957-6.

Amer-Yahia:2006:DXT

[AYBH^{+}06] Sihem Amer-Yahia, Zohra Bellahsène, Ela Hunt, Rainer Unland, and Jeffrey Xu Yu, editors. Database and XML Technologies: 4th International XML Database Symposium, XSym 2006 Seoul, Korea, September 10–11, 2006 Proceedings, volume 4156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38877-X (print), 3-540-38879-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38879-1.

Aykin:2007:UIG

[Ayk07a] Nuray Aykin, editor. Usability and Internationalization. Global and Local User Interfaces: Second International Conference on Usability and Internationalization, UI-HCII 2007, Held as Part of HCI International 2007, Beijing, China, July 22– 27, 2007, Proceedings, Part II, volume 4560 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73288-8 (print), 3-540-73289-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73289-1.

Aykin:2007:UIH

[Ayk07b] Nuray Aykin, editor. Usability and Internationalization. HCI and Culture: Second International Conference on Usability and Internationalization, UI-HCII 2007, Held as Part of HCI International 2007, Beijing, China, July 22– 27, 2007, Proceedings, Part I, volume 4559 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73286-1 (print), 3-540-73287-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73287-7.

Aykin:2009:IDG

[Ayk09] Nuray Aykin, editor. Internationalization, Design and Global Development: Third International Conference, IDGD 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009. Proceedings, volume 5623 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02766-0 (print), 3-642-02767-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02767-3.

Aguilera:2011:DCN

[AYV^{+}11] Marcos K. Aguilera, Haifeng Yu, Nitin H. Vaidya, Vikram Srinivasan, and Romit Roy Choudhury, editors. Distributed Computing and Networking: 12th International Conference, ICDCN 2011, Bangalore, India, January 2– 5, 2011. Proceedings, volume 6522 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17678-X (print), 3-642-17679-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17679-1.

Adorni:1996:TNL

[AZ96] Giovanni Adorni and Michael Zock, editors. Trends in natural language generation: an artificial intelligence perspective: fourth European workshop, EWNLG ’93, Pisa, Italy, April 28–30, 1993: Selected Papers, volume 1036 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60800-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 E89 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1036.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60800-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1036.

Ardagna:2011:IST

[AZ11] Claudio A. Ardagna and Jianying Zhou, editors. Information Security Theory and Practice. Security and Privacy of Mobile Devices in Wireless Communication: 5th IFIP WG 11.2 International Workshop, WISTP 2011, Heraklion, Crete, Greece, June 1–3, 2011. Proceedings, volume 6633 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21039-2 (print), 3-642-21040-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21040-2.

Barros:2012:AAI

[B^{+}12a] Leliane N. Barros et al., editors. Advances in artificial intelligence — SBIA 2012: 21th [i.e.: 21st] Brazilian Symposium on Artificial Intelligence, Curitiba, Brazil, October 20–25, 2012, Proceedings, volume 7589 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34458-5, 3-642-34459-3 (e-book). LCCN Q334 .S56 2012. URL http:

//www.springerlink.
com/content/978-3-642-
34458-9;   http://www.

                                        springerlink.com/content/
                                        978-3-642-34459-6.

Bshouty:2012:ALT

[B^{+}12b] Nader H. Bshouty et al., editors. Algorithmic learning theory: 23rd International Conference, ALT 2012, Lyon, France, October 29–31, 2012. Proceedings, volume 7568 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34105-5, 3-642-34106-3 (e-book). LCCN QA76.9.A43 A48 2012.

                                        URL http://www.springerlink.
                                        com/content/978-3-642-
                                        34105-2;   http://www.
                                        springerlink.com/content/
                                        978-3-642-34106-9.

Barahona:1999:PAI

[BA99] P. Barahona and Jose Julio Alferes, editors. Progress in artificial intelligence: 9th Portuguese Conference on Artificial Intelligence, EPIA’99, Evora, Portugal, September 21– 24, 1999: Proceedings, volume 1695 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66548-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA267.A1 L43 no.1695. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1695.htm;
http://www.springerlink.
com/content/978-3-540-
66548-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1695.

Bagchi:2006:ISS

[BA06] Aditya Bagchi and Vijayalakshmi Atluri, editors. Information Systems Security: Second International Conference, ICISS 2006, Kolkata, India, December 19–21, 2006. Proceedings, volume 4332 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68962-1 (print), 3-540-68963-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68963-8.

Barker:2007:DAS

[BA07] Steve Barker and Gail-Joon Ahn, editors. Data and Applications Security XXI: 21st Annual IFIP WG 11.3 Working Conference on Data and Applications Security, Redondo Beach, CA, USA, July 8–11, 2007. Proceedings, volume 4602 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73533-X (print), 3-540-73538-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73538-0.

Baader:2003:ADC

[Baa03] Franz Baader, editor. Automated Deduction: CADE-19: 19th International Conference on Automated Deduction, Miami Beach, FL, USA, July 28–August 2, 2003: Proceedings, volume 2741 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40559-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I57 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2741.
htm;   http://www.springerlink.
com/content/978-3-540-
40559-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2741;
 http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b11829.

Baader:2007:TRA

[Baa07] Franz Baader, editor.

Term Rewriting and Applications: 18th International Conference, RTA 2007, Paris, France, June 26–28, 2007. Proceedings, volume 4533 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73447-3 (print), 3-540-73449-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73449-9.

Bachmair:2000:RTA

[Bac00] Leo Bachmair, editor.

Rewriting techniques and applications: 11th International Conference, RTA 2000, Norwich, UK, July 10–12, 2000: Proceedings, volume 1833 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67778-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1833.

URL http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t1833.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67778-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1833.

Briggs:2008:GDI

[BAdVR08] Robert O. Briggs, Pedro Antunes, Gert-Jan de Vreede, and Aaron S. Read, editors. Groupware: Design, Implementation, and Use: 14th International Workshop, CRIWG 2008, Omaha, NE, USA, September 14–18, 2008, Revised Selected Papers, volume 5411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92830-8 (print), 3-540-92831-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        92831-7.

Bras-Amoros:2009:AAA

[BAH09] Maria Bras-Amorós and Tom Høholdt, editors. Applied Algebra, Algebraic Algorithms and Error-Correcting Codes: 18th International Symposium, AAECC-18 2009, Tarragona, Spain, June 8–12, 2009. Proceedings, volume 5527 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02180-8 (print), 3-642-02181-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02181-7.

Bellazzi:2007:AIM

[BAHH07] Riccardo Bellazzi, Ameen Abu-Hanna, and Jim Hunter, editors. Artificial Intelligence in Medicine: 11th Conference on Artificial Intelligence in Medicine, AIME 2007, Amsterdam, The Netherlands, July 7–11, 2007. Proceedings, volume 4594 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73598-4 (print), 3-540-73599-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73599-1.

Baik:2005:SMS

[Bai05] Doo-Kwon Baik, editor. Systems modeling and simulation: theory and applications: Third Asian Simulation Conference, AsiaSim 2004, Jeju Island, Korea, October 4–6, 2004: Revised selected papers, volume 3398 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24477-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65. URL

                                        http://www.springerlink.
                                        com/content/978-3-540-
                                        24477-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3398;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105611.

Baker:1995:MMI

[Bak95] Henry G. Baker, editor. Memory management: International Workshop, IWMM 95, Kinross, UK, September 27–29, 1995: proceedings, volume 986 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60368-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M45 I96 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0986.
htm;   http://www.springerlink.
com/content/978-3-540-
60368-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=986.

Boscarol:1988:FLF

[BAL88] M. Boscarol, L. Carlucci Aiello, and G. Levi, editors. Foundations of Logic and Functional Programming Workshop: Trento, Italy, December 15–19, 1986: proceedings, volume 306 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19129-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .F68581 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0306.htm;
 http://www.springerlink.
com/content/978-0-387-
19129-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=306.

Banerjee:1992:LCP

[Ban92] Utpal Banerjee, editor. Languages and compilers for parallel computing: fourth International Workshop, Santa Clara, California, USA, August 7–9, 1991: proceedings, volume 589 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55422-X (Berlin), 0-387-55422-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .L36 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0589.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55422-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=589.

Banerjee:1993:LCP

[Ban93] Utpal Banerjee, editor. Languages and compilers for parallel computing: 5th International Workshop, New Haven, Connecticut, USA, August 3–5, 1992: proceedings, volume 757 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57502-2, 0-387-57502-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .L36 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0757.htm;
 http://www.springerlink.
com/content/978-3-540-
57502-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=757.

Banerjee:1994:LCP

[Ban94] Utpal Banerjee, editor.

Languages and compilers for parallel computing: 6th International Workshop, Portland, Oregon, USA, August 12–14, 1993: proceedings, volume 768 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57659-2 (Berlin), 0-387-57659-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1993. DM122.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0768.
htm;   http://www.springerlink.
com/content/978-3-540-
57659-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=768.

Banzhaf:1998:GPF

[Ban98] Wolfgang Banzhaf, editor. Genetic programming: first European workshop, EuroGP’98: Paris, France, April 14–15, 1998: Proceedings, volume 1391 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64360-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.623 .E97 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1391.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64360-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1391.

Boutaba:2005:NNT

[BAP^{+}05] Raouf Boutaba, Kevin Almeroth, Ramon Puigjaner, Sherman Shen, and James P. Black, editors. NETWORK-ING 2005: Networking Technologies, Services, and Protocols; Performance of Computer and Communication Networks; Mobile and Wireless Communication Systems. 4th International IFIP-TC6 Networking Conference Waterloo, Canada, May 2–6, 2005 Proceedings, volume 3462 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25809-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25809-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3462;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136094.

Barringer:1985:SVT

[Bar85] Howard Barringer. A survey of verification techniques for parallel programs, volume 191 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15239-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 114 pp. LCCN QA76.5.B29627 1985. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0191.htm;

                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        15239-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=191.

Barnes:1997:ARL

[Bar97] John G. P. Barnes. Ada 95 rationale: the language, the standard libraries, volume 1247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63143-7. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 458 pp. LCCN QA76.73.A35 A217 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1247.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63143-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1247.

Document to accompany the new Ada standard, officially ISO/IEC 8652: 1995.

Barkowsky:2002:MRP

[Bar02] Thomas Barkowsky. Mental Representation and Processing of Geographic Knowledge: A Computational Approach, volume 2541 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00216-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 174 pp. LCCN G71.5 .B37 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2541.htm;
 http://www.springerlink.
com/content/978-3-540-
00216-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2541.

Based on author’s doctoral dissertation, Department of Informatics, University of Hamburg.

Bartenstein:2003:WKM

[Bar03] Oskar Bartenstein, editor. Web knowledge management and decision support: 14th International Conference on Applications of Prolog, INAP 2001, Tokyo, Japan, October 20–22, 2001: Revised Papers, volume 2543 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00680-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.P76 .I58 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2543.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00680-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2543.

Barbero:2008:CTA

[Bar08a] Ángela Barbero, editor. Coding Theory and Applications: Second International Castle Meeting, ICMCTA 2008, Castillo de la Mota, Medina del Campo, Spain, September 15–19, 2008. Proceedings, volume 5228 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87447-X (print), 3-540-87448-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87448-5.

Barnes:2008:ARL

[Bar08b] John Barnes, editor. Ada 2005 Rationale: The Language, The Standard Libraries, volume 5020 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CO-DEN LNCSD9. ISBN 3-540-79700-9 (print), 3-540-79701-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 259 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79701-2.

Barthe:2011:PLS

[Bar11] Gilles Barthe, editor.

Programming Languages and Systems: 20th European Symposium on Programming, ESOP 2011, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2011, Saarbrücken, Germany, March 26–April 3, 2011. Proceedings, volume 6602 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19717-5 (print), 3-642-19718-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19718-5.

Batini:2001:CIS

[Bat01] Carlo Batini, editor.

Cooperative information systems: 9th International Conference, CoopIS 2001, Trento, Italy, September 5–7, 2001: Proceedings, volume 2172 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42524-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I158 2001; QA267.A1 L43 no.2172.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2172.
htm;   http://www.springerlink.
com/content/978-3-540-
42524-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2172.

Batzoglou:2009:RCM

[Bat09] Serafim Batzoglou, editor. Research in Computational Molecular Biology: 13th Annual International Conference, RECOMB 2009, Tucson, AZ, USA, May 18–21, 2009. Proceedings, volume 5541 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02007-0 (print), 3-642-02008-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02008-7.

Bauer:1975:SEA

[Bau75] Friedrich Ludwig Bauer.

Software engineering: an advanced course, volume 30 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07168-7. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 545 pp. LCCN QA76.6 .A331 1972a. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0030.htm;
 http://www.springerlink.
com/content/978-0-387-
07168-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=30.

“The advanced course took place February 21– March 3, 1972, organized by the Mathematical Institute of the Technical University of Munich and the Leibni[t]z Computing Center of the Bavarian Academy of Sciences, in cooperation with the European Communities, sponsored by the Ministry of Education and Science of the Federal Republic of Germany.”

First published in 1973 under title: Advanced Course on Software Engineering.

Baumgartner:1998:TRC

[Bau98] Peter Baumgartner. Theory reasoning in connection calculi, volume 1527 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65509-3. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 283 pp. LCCN QA76.9.A96 B38 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1527.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65509-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1527.

Baumgart:1999:SNC

[Bau99] Rainer Baumgart, editor. Secure networking — CQRE [Secure] ’99: International Exhibition and Congress, Düsseldorf, Germany, November 30–December 2, 1999: Proceedings, volume 1740 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66800-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.59 .C67 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1740.htm;
http://www.springerlink.
com/content/978-3-540-
66800-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1740.

Baumann:2000:MAC

[Bau00] Joachim Baumann. Mobile agents: control algorithms, volume 1658 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41192-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 161 pp. LCCN QA76.76.I58 B38 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1658.
htm;   http://www.springerlink.
com/content/978-3-540-
41192-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1658.

Bubak:2000:HPC

[BAWH00] Marian Bubak, Hamideh Afsarmanesh, Roy Williams, and Bob Hertzberger, editors. High performance computing and networking: 8th International Conference, HPCN Europe 2000, Amsterdam, The Netherlands, May 8–10, 2000: Proceedings, volume 1823 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67553-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1823. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1823.htm;
 http://www.springerlink.
com/content/978-3-540-
67553-2;   http://www.
springerlink.com/link.
asp?id=fundvl8nd68e;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=1823.

Bauer:1979:PCI

[BB79] Friedrich Ludwig Bauer and M. Broy, editors. Program construction: international summer school, volume 69 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CODEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0069.htm;
 http://www.springerlink.
com/content/978-0-387-
09251-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=69.

“The international summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Beth:1989:LAN

[BB89] Thomas Beth and T. A.

Berson, editors. Local area network security: Workshop LANSEC ’89, European Institute for System Security (E.I.S.S.), Karlsruhe, FRG, April 3–6, 1989: proceedings, volume 396 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51754-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.7 .W65 1989. US$17.80. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0396.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51754-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=396.

Barter:1990:AAJ

[BB90] C. J. Barter and Michael J. Brooks, editors. AI ’88: 2nd Australian Joint Artificial Intelligence Conference, Adelaide, Australia, November 15–18, 1988: proceedings, volume 406 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-52062-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.406. DM63.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0406.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        52062-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=406.

Barzdin:1991:BCS

[BB91] Ia. M. Barzdin’ and D. Bjorner. Baltic computer science: Selected Papers, volume 502 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54131-4 (Berlin), 0-387-54131-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 619 pp. LCCN QA76 .B26 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0502.htm;
 http://www.springerlink.
com/content/978-3-540-
54131-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=502.

Beilner:1995:QEC

[BB95] Heinz Beilner and Falko Bause, editors. Quantitative evaluation of computing and communication systems: 8th International Conference on Modelling Techniques and Tools for Computer Performance Evaluation, Performance Tools ’95, 8th GI/ITG Conference on Measuring, Modelling, and Evaluating Computing and Communication Systems, MMB ’95, Heidelberg, Germany, September 20– 22, 1995: proceedings, volume 977 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60300-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.E94 I55 1995.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0977.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60300-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=977.

Berquist:1996:MIH

[BB96] Kim Berquist and Andrew Berquist, editors. Managing information highways: the PRISM book: principles, methods, and case studies for designing telecommunications management systems, volume 1164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-62008-7. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 417 pp. LCCN TK5103.7.M35 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1164.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62008-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1164.

Beeri:1999:DTI

[BB99] Catriel Beeri and Peter Buneman, editors. Database theory: ICDT’99: 7th International Conference, Jerusalem, Israel, January 10–12, 1999: Proceedings, volume 1540 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65452-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I557 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1540.
htm;   http://www.springerlink.
com/content/978-3-540-
65452-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1540.

Borbinha:2000:RAT

[BB00a] Jose Borbinha and Thomas Baker, editors. Research and Advanced Technology for Digital Libraries: 4th European Conference, ECDL 2000, Lisbon, Portugal, September 18–20, 2000. Proceedings, volume 1923 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41023-6 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1923. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1923.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41023-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1923.

Buyya:2000:GCG

[BB00b] Rajkumar Buyya and Mark Baker, editors. Grid Computing—GRID 2000: First IEEE/ACM International Workshop, Bangalore, India, December 17, 2000. Proceedings, volume 1971 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41403-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1971. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1971.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41403-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1971.

Benferhat:2001:SQA

[BB01a] Salem Benferhat and Philippe Besnard, editors. Symbolic and quantitative approaches to reasoning with uncertainty: 6th European Conference, ECSQARU 2001, Toulouse, France, September 19–21, 2001: Proceedings, volume 2143 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42464-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q375 .E87 2001; QA267.A1 L43 no.2143. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2143.htm;
http://www.springerlink.
com/content/978-3-540-
42464-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2143.

Beun:2001:CMC

[BB01b] Robbert-Jan Beun and Harry C. Bunt, editors. Cooperative multimodal communication: Second International Conference, CMC’98, Tilburg, The Netherlands, January 28–30, 1998: Selected Papers, volume 2155 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42806-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I596 1998; QA267.A1 L43 no.2155. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2155.htm;
 http://www.springerlink.
com/content/978-3-540-
42806-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2155.

Bruel:2002:AOO

[BB02] Jean-Michel Bruel and Zohra Bellahsène, editors. Advances in Object-Oriented Information Systems: OOIS 2002 Workshops, Montpellier, France, September 2, 2002. Proceedings, volume 2426 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44088-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55875 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2426.htm;
 http://www.springerlink.
com/content/978-3-540-
44088-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2426.

Also available via the World Wide Web.

Bianchi-Berthouze:2003:DNI

[BB03] Nadia Bianchi-Berthouze, editor. Databases in Networked Information Systems: Third International Workshop, DNIS 2003, Aizu, Japan, September 22–24, 2003: Proceedings, volume 2822 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20111-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D5862 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2822.
htm;   http://www.springerlink.
com/content/978-3-540-
20111-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2822;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13478.

Bengio:2005:MLM

[BB05a] Samy Bengio and Hervé Bourlard, editors. Machine learning for multimodal interaction: first international workshop, MLMI 2004, Martigny, Switzerland, June 21–23, 2004. Revised selected papers, volume 3361 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24509-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .W69 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24509-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3361;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105752.

Bernardo:2005:FMM

[BB05b] Marco Bernardo and Alessandro Bogliolo, editors. Formal methods for mobile computing: 5th international school on formal methods for the design of computer, communication, and software systems, SFM-MOBY 2005, Bertinoro, Italy, April 26–30, 2005, advanced lectures, volume 3465 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25697-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25697-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3465;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b135888.

Bonchi:2006:KDI

[BB06] Francesco Bonchi and Jean-François Boulicaut, editors. Knowledge Discovery in Inductive Databases: 4th International Workshop, KDID 2005, Porto, Portugal, October 3, 2005, Revised Selected and Invited Papers, volume 3933 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33292-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-

                                        33292-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3933.

Bandara:2007:IDM

[BB07] Arosha K. Bandara and Mark Burgess, editors. Inter-Domain Management: First International Conference on Autonomous Infrastructure, Management and Security, AIMS 2007, Oslo, Norway, June 21– 22, 2007. Proceedings, volume 4543 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72985-2 (print), 3-540-72986-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72986-0.

Blum:2010:LIO

[BB10] Christian Blum and Roberto Battiti, editors. Learning and Intelligent Optimization: 4th International Conference, LION 4, Venice, Italy, January 18–22, 2010. Selected Papers, volume 6073 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13799-7 (print), 3-642-13800-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13800-3.

Bravetti:2011:WSF

[BB11] Mario Bravetti and Tevfik Bultan, editors. Web Services and Formal Methods: 7th International Workshop, WS-FM 2010, Hoboken, NJ, USA, September 16–17, 2010. Revised Selected Papers, volume 6551 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19588-1 (print), 3-642-19589-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19589-1.

Barneva:2012:CIA

[BBA12] Reneta P. Barneva, Valentin E. Brimkov, and Jake K. Aggarwal, editors. Combinatorial Image Analaysis: 15th International Workshop, IWCIA 2012, Austin, TX, USA, November 28– 30, 2012. Proceedings, volume 7655 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34731-2 (print), 3-642-34732-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34732-0.

Bartz-Beielstein:2007:HMI

[BBAB^{+}07] Thomas Bartz-Beielstein, María José Blesa Aguilera, Christian Blum, Boris Naujoks, Andrea Roli, Günter Rudolph, and Michael Sampels, editors. Hybrid Metaheuristics: 4th International Workshop, HM 2007, Dortmund, Germany, October 8–9, 2007. Proceedings, volume 4771 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75513-6 (print), 3-540-75514-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75514-2.

Bauer:1985:MPCa

[BBB^{+}85] Friedrich Ludwig Bauer, R. Berghammer, M. Broy, W. Dosch, F. Geiselbrechtinger, R. Gnatz, E. Hangel, W. Hesse, B. Krieg-Brückner, A. Laut, T. Matzner, B. Möller, F. Nickl, H. Partsch, R. Pepper, K. Samelson, M. Wirsing, and H. Wössner, editors. The Munich Project CIP: The Wide Spectrum Language CIP-L, volume 183 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15187-7 (New York), 3-540-15187-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 275 pp. LCCN QA76.6 .M86 1985 v. 1–2 (c1985– 1987). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0183.htm;
http://www.springerlink.
com/content/978-0-387-
15187-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=183.

Bunt:1998:MHC

[BBB98] Harry C. Bunt, Robbert-Jan Beun, and Tijn Borghuis, editors. Multimodal human-computer communication: systems, techniques, and experiments, volume 1374 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64380-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85M84 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1374.
htm;   http://www.springerlink.
com/content/978-3-540-
64380-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1374.

Bui:2006:IIC

[BBB06] Alain Bui, Marc Bui, and Thomas Böhme, editors. Innovative Internet Community Systems: 5th International Workshop, IICS 2005, Paris, France, June 20– 22, 2005. Revised Papers, volume 3908 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33973-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-

33973-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3908.

Barbosa:2007:DXT

[BBB^{+}07a] Denilson Barbosa, Angela Bonifati, Zohra Bellahsène, Ela Hunt, and Rainer Unland, editors. Database and XML Technologies: 5th International XML Database Symposium, XSym 2007, Vienna, Austria, September 23–24, 2007. Proceedings, volume 4704 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75287-0 (print), 3-540-75288-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75288-2.

Bemporad:2007:HSC

[BBB07b] Alberto Bemporad, Antonio Bicchi, and Giorgio Buttazzo, editors. Hybrid Systems: Computation and Control: 10th International Workshop, HSCC 2007, Pisa, Italy, April 3–5, 2007. Proceedings, volume 4416 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71492-8 (print), 3-540-71493-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71493-4.

Borger:2008:ASM

[BBBB08] Egon Börger, Michael Butler, Jonathan P. Bowen, and Paul Boca, editors. Abstract State Machines, B and Z: First International Conference, ABZ 2008, London, UK, September 16– 18, 2008. Proceedings, volume 5238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87602-2 (print), 3-540-87603-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87603-8.

Buschmann:2004:OOT

[BBC04] Frank Buschmann, Alejandro P. Buchmann, and Mariano A. Cilia, editors. Object-Oriented Technology: ECOOP 2003 workshop reader: ECOOP 2003 workshops, Darmstadt, Germany, July 21–25, 2003: final reports, volume 3013 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22405-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35 E33 2003b.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3013.
htm;   http://www.springerlink.
com/content/978-3-540-
22405-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3013;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98806.

Blesa:2008:HMI

[BBC^{+}08] María J. Blesa, Christian Blum, Carlos Cotta, Antonio J. Fernández, José E. Gallardo, Andrea Roli, and Michael Sampels, editors. Hybrid Metaheuristics: 5th International Workshop, HM 2008, Málaga, Spain, October 8–9, 2008. Proceedings, volume 5296 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88438-6 (print), 3-540-88439-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88439-2.

Baek:2008:PSS

[BBCL08] Joonsang Baek, Feng Bao, Kefei Chen, and Xuejia Lai, editors. Provable Security: Second International Conference, ProvSec 2008, Shanghai, China, October 30 — November 1, 2008. Proceedings, volume 5324 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88732-6 (print), 3-540-88733-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88733-1.

Bongiovanni:1997:ACT

[BBD97] Giancarlo Bongiovanni, D. Bovet, and Giuseppe Di Battista, editors. Algorithms and complexity: third Italian conference, CIAC ’97, Rome, Italy, March 12–14, 1997: proceedings, volume 1203 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62592-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .I885 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1203.htm;
 http://www.springerlink.
com/content/978-3-540-
62592-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1203.

Papers presented at the Third Conference on Algorithms and Complexity, Mar. 12–14, 1997 in Rome.

Blesa:2009:HMI

[BBD^{+}09] María J. Blesa, Christian Blum, Luca Di Gaspero, Andrea Roli, Michael Sampels, and Andrea Schaerf, editors. Hybrid Metaheuristics: 6th International Workshop, HM 2009, Udine, Italy, October 16–17, 2009. Proceedings, volume 5818 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04917-6 (print), 3-642-04918-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        04918-7.

Bacardit:2010:LCS

[BBD^{+}10] Jaume Bacardit, Will Browne, Jan Drugowitsch, Ester Bernadó-Mansilla, and Martin V. Butz, editors. Learning Classifier Systems: 11th International Workshop, IWLCS 2008, Atlanta, GA, USA, July 13, 2008, and 12th International Workshop, IWLCS 2009, Montreal, QC, Canada, July 9, 2009, Revised Selected Papers, volume 6471 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17507-4 (print), 3-642-17508-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17508-4.

Baeten:2012:TCS

[BBdB12] Jos C. M. Baeten, Tom Ball, and Frank S. de Boer, editors. Theoretical Computer Science: 7th IFIP TC 1/ WG 2.2 International Conference, TCS 2012, Amsterdam, The Netherlands, September 26– 28, 2012. Proceedings, volume 7604 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33474-1 (print), 3-642-33475-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33475-7.

Beer:2012:AEG

[BBDS12] Martin Beer, Cyril Brom, Frank Dignum, and Von-Wun Soo, editors. Agents for Educational Games and Simulations: International Workshop, AEGS 2011, Taipei, Taiwan, May 2, 2011. Revised Papers, volume 7471 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32325-1 (print), 3-642-32326-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32326-3.

Baader:2001:KAA

[BBE01] Franz Baader, Gerhard Brewka, and Thomas Eiter, editors. KI 2001: advances in artificial intelligence: Joint German/Austrian Conference on AI, Vienna, Austria, September 19– 21, 2001: Proceedings, volume 2174 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42612-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2174. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2174.htm;
 http://www.springerlink.
com/content/978-3-540-
42612-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2174.

Barth:1998:DLC

[BBEdK98] Andreas Barth, Michael Breu, Albert Endres, and Arnoud de Kemp, editors. Digital libraries in computer science: the MeDoc approach, volume 1392 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64493-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 238 pp. LCCN ZA4082.G33 .M434 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1392.htm;
http://www.springerlink.
com/content/978-3-540-
64493-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1392.

Barahona:2006:RWS

[BBF^{+}06] Pedro Barahona, François Bry, Enrico Franconi, Nicola Henze, and Ulrike Sattler, editors. Reasoning Web: Second International Summer School 2006, Lisbon, Portugal, September 4–8, 2006, Tutorial Lectures, volume 4126 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38409-X (print), 3-540-38412-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38412-0.

Becker:2003:DRM

[BBGR03] Eberhard Becker, Willms Buhse, Dirk Günnewig, and Niels Rump, editors. Digital Rights Management: Technological, Economic, Legal and Political Aspects, volume 2770 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40465-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 805 pp. LCCN QA76.9.A25 D53 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2770.htm;
 http://www.springerlink.
com/content/978-3-540-
40465-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2770;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12637.

Balcazar:2010:MLKa

[BBGS10a] José Luis Balcázar, Francesco Bonchi, Aristides Gionis, and Michèle Sebag, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2010, Barcelona, Spain, September 20–24, 2010, Proceedings, Part I, volume 6321 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15879-X (print), 3-642-15880-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15880-3.

Balcazar:2010:MLKb

[BBGS10b] José Luis Balcázar, Francesco Bonchi, Aristides Gionis, and Michèle Sebag, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2010, Barcelona, Spain, September 20–24, 2010, Proceedings, Part II, volume 6322 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15882-X (print), 3-642-15883-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15883-4.

Balcazar:2010:MLKc

[BBGS10c] José Luis Balcázar, Francesco Bonchi, Aristides Gionis, and Michèle Sebag, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2010, Barcelona, Spain, September 20–24, 2010, Proceedings, Part III, volume 6323 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15938-9 (print), 3-642-15939-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15939-8.

Blanck:2001:CCA

[BBH01] Jens Blanck, Vasco Brattka, and Peter Hertling, editors. Computability and complexity in analysis: 4th International Workshop, CCA 2000, Swansea, UK, September 17–19, 2000, Selected Papers, volume 2064 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42197-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA299.6 .C6585 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2064.htm;
 http://www.springerlink.

com/content/978-3-540-
42197-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2064.

Barthe:2005:CAS

[BBH^{+}05] Gilles Barthe, Lilian Burdy, Marieke Huisman, Jean-Louis Lanet, and Traian Muntean, editors. Construction and analysis of safe, secure, and interoperable smart devices: international workshop, CAS-SIS 2004, Marseille, France, March 10–14, 2004, Revised selected papers, volume 3362 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24287-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.S62 C38 2004. URL http:

//www.springerlink.
com/content/978-3-540-
24287-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3362;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105030.

Brimkov:2008:CIA

[BBH08] Valentin E. Brimkov, Reneta P. Barneva, and Herbert A. Hauptman, editors. Combinatorial Image Analysis: 12th International Workshop, IWCIA 2008, Buffalo, NY, USA, April 7– 9, 2008. Proceedings, volume 4958 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78274-5 (print), 3-540-78275-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78275-9.

Barneva:2010:CMO

[BBH^{+}10] Reneta P. Barneva, Valentin E. Brimkov, Herbert A. Hauptman, Renato M. Natal Jorge, and João Manuel R. S. Tavares, editors. Computational Modeling of Objects Represented in Images: Second International Symposium, CompIMAGE 2010, Buffalo, NY, USA, May 5–7, 2010. Proceedings, volume 6026 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12711-8 (print), 3-642-12712-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12712-0.

Breslin:2011:RTD

[BBK^{+}11] John G. Breslin, Thomas N.

Burg, Hong-Gee Kim, Tom Raftery, and Jan-Hinrik Schmidt, editors. Recent Trends and Developments in Social Software: International Conferences on Social Software, BlogTalk 2008, Cork, Ireland, March 3– 4, 2008, and BlogTalk 2009, Jeju Island, South Korea, September 15–16, 2009. Revised Selected Papers, volume 6045 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16580-X (print), 3-642-16581-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16581-8.

Becvar:2012:NWI

[BBK12] Zdenek Becvar, Robert Bestak, and Lukas Kencl, editors. NETWORK-ING 2012 Workshops: International IFIP TC 6 Workshops, ETICS, HetsNets, and CompNets, Held at NETWORKING 2012, Prague, Czech Republic, May 25, 2012. Proceedings, volume 7291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30038-3 (print), 3-642-30039-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30039-4.

Biswas:2011:AIE

[BBKM11] Gautam Biswas, Susan Bull, Judy Kay, and Antonija Mitrovic, editors. Artificial Intelligence in Education: 15th International Conference, AIED 2011, Auckland, New Zealand, June 28 — July 2011, volume 6738 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21868-7 (print), 3-642-21869-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21869-9.

Bebis:2006:AVCa

[BBKP06] George Bebis, Richard Boyle, Darko Koracin, and Bahram Parvin, editors. Advances in visual computing: first international symposium, ISVC 2005, Lake Tahoe, NV, USA, December 5– 7, 2005. proceedings, volume 3804 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30750-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30750-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3804.

Bert:2003:ZFS

[BBKW03] Didier Bert, Jonathan P.

Bowen, Steve King, and Marina Waldén, editors. ZB 2003: formal specification and development in Z and B: Third International Conference of B and Z Users, Turku, Finland, June 4–6, 2003: Proceedings, volume 2651 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40253-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.B155 I58 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2651.htm;
 http://www.springerlink.
com/content/978-3-540-
40253-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2651.

Also available via the World Wide Web.

Budach:1987:FCT

[BBL87] L. Budach, Rais Gatich Bukharaev, and O. B. Lupanov, editors. Fundamentals of computation theory: International Conference, FCT’87, Kazan, USSR, June 22–26, 1987: proceedings, volume 278 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18740-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I6141 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0278.htm;
 http://www.springerlink.
com/content/978-0-387-
18740-2;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=278.

Beckmann:2006:LAC

[BBLT06] Arnold Beckmann, Ulrich Berger, Benedikt Löwe, and John V. Tucker, editors. Logical Approaches to Computational Barriers: Second Conference on Computability in Europe, CiE 2006, Swansea, UK, June 30–July 5, 2006. Proceedings, volume 3988 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35466-2 (print), 3-540-35468-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35468-0.

Baroglio:2008:RWI

[BBM^{+}08] Cristina Baroglio, Piero A.

Bonatti, Jan Małuszyński, Massimo Marchiori, Axel Polleres, and Sebastian Schaffert, editors. Reasoning Web: 4th International Summer School 2008, Venice, Italy, September 7–11, 2008, Tutorial Lectures, volume 5224 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85656-0 (print), 3-540-85658-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85658-0.

Bacardit:2008:LCS

[BBMB^{+}08] Jaume Bacardit, Ester Bernadó-Mansilla, Martin V. Butz, Tim Kovacs, Xavier Llorà, and Keiki Takadama, editors. Learning Classifier Systems: 10th International Workshop, IWLCS 2006, Seattle, MA, USA, July 8, 2006 and 11th International Workshop, IWLCS 2007, London, UK, July 8, 2007, Revised Selected Papers, volume 4998 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88137-9 (print), 3-540-88138-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88138-4.

Boughanem:2009:AIR

[BBMSD09] Mohand Boughanem, Catherine Berrut, Josiane Mothe, and Chantal Soule-Dupuy, editors. Advances in Information Retrieval: 31th European Conference on IR Research, ECIR 2009, Toulouse, France, April 6–9, 2009. Proceedings, volume 5478 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00957-3 (print), 3-642-00958-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00958-7.

Banieqbal:1989:TLS

[BBP89] B. Banieqbal, Howard Barringer, and A. Pnueli, editors. Temporal logic in specification: Altrincham, UK, April 8–10, 1987: proceedings, volume 398 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51803-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T42 1989. US$34.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0398.htm;

                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51803-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=398.

Based on papers presented at the Colloquium on Temporal Logic in Specification.

Bjorner:1993:FMP

[BBP93] D. Bjorner, M. Broy, and I. V. Pottosin, editors. Formal methods in programming and their applications: international conference, Academgorodok, Russia, June/July 1993: proceedings, volume 735 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57316-X (USA), 3-540-57316-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .F577 1993. DM80.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0735.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57316-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=735.

Bjorner:1996:PSI

[BBP96] D. (Dines) Bjorner, M. Broy, and I. V. Pottosin, editors. Perspectives of system informatics: Second International Andrei Ershov Memorial Conference, Akademgorodok, Novosibirsk, Russia, June 1996: proceedings, volume 1181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-62064-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .A454 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1181.htm;
 http://www.springerlink.
com/content/978-3-540-
62064-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1181.

Bauknecht:2004:CWT

[BBP04] Kurt Bauknecht, Martin Bichler, and Birgit Pröll, editors. E-commerce and Web Technologies: 5th International Conference, EC-Web 2004, Zaragoza, Spain, August 31–September 3, 2004. Proceedings, volume 3182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22917-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.32 .I5796 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        22917-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3182;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100074.

Bebis:2006:AVCb

[BBP^{+}06a] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Paolo Remagnino, Ara Nefian, Gopi Meenakshisundaram, Valerio Pascucci, Jiri Zara, Jose Molineros, Holger Theisel, and Tom Malzbender, editors. Advances in Visual Computing: Second International Symposium, ISVC 2006 Lake Tahoe, NV, USA, November 6–8, 2006 Proceedings, Part I, volume 4291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48628-3 (print), 3-540-48631-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
48631-2.

Bebis:2006:AVCc

[BBP^{+}06b] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Paolo Remagnino, Ara Nefian, Gopi Meenakshisundaram, Valerio Pascucci, Jiri Zara, Jose Molineros, Holger Theisel, and Tom Malzbender, editors. Advances in Visual Computing: Second International Symposium, ISVC 2006 Lake Tahoe, NV, USA, November 6–8, 2006. Proceedings, Part II, volume 4292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48626-7 (print), 3-540-48627-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48627-5.

Bebis:2007:AVCa

[BBP^{+}07a] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Nikos Paragios, Syeda-Mahmood Tanveer, Tao Ju, Zicheng Liu, Sabine Coquillart, Carolina Cruz-Neira, Torsten Müller, and Tom Malzbender, editors. Advances in Visual Computing: Third International Symposium, ISVC 2007, Lake Tahoe, NV, USA, November 26–28, 2007, Proceedings, Part I, volume 4841 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76857-2 (print), 3-540-76858-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76858-6.

Bebis:2007:AVCb

[BBP^{+}07b] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Nikos Paragios, Syeda-Mahmood Tanveer, Tao Ju, Zicheng Liu, Sabine Coquillart, Carolina Cruz-Neira, Torsten Müller, and Tom Malzbender, editors. Advances in Visual Computing: Third International Symposium, ISVC 2007, Lake Tahoe, NV, USA, November 26–28, 2007, Proceedings, Part II, volume 4842 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76855-6 (print), 3-540-76856-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76856-2.

Bebis:2008:AVCa

[BBP^{+}08a] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Paolo Remagnino, Fatih Porikli, Jörg Peters, James Klosowski, Laura Arns, Yu Ka Chun, Theresa-Marie Rhyne, and Laura Monroe, editors. Advances in Visual Computing: 4th International Symposium, ISVC 2008, Las Vegas, NV, USA, December 1–3, 2008. Proceedings, Part I, volume 5358 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89638-4 (print), 3-540-89639-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89639-5.

Bebis:2008:AVCb

[BBP^{+}08b] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Paolo Remagnino, Fatih Porikli, Jörg Peters, James Klosowski, Laura Arns, Yu Ka Chun, Theresa-Marie Rhyne, and Laura Monroe, editors. Advances in Visual Computing: 4th International Symposium, ISVC 2008, Las Vegas, NV, USA, December 1–3, 2008. Proceedings, Part II, volume 5359 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89645-7 (print), 3-540-89646-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89646-3.

Bebis:2009:AVCb

[BBP^{+}09a] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Yoshinori Kuno, Junxian Wang, Renato Pajarola, Peter Lindstrom, André Hinkenjann, Miguel L. Encarnação, Cláudio T. Silva, and Daniel Coming, editors. Advances in Visual Computing: 5th International Symposium, ISVC 2009, Las Vegas, NV, USA, November 30–December 2, 2009. Proceedings, Part II, volume 5876 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10519-X (print), 3-642-10520-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10520-3.

Bebis:2009:AVCa

[BBP^{+}09b] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Yoshinori Kuno, Junxian Wang, Jun-Xuan Wang, Junxian Wang, Renato Pajarola, Peter Lindstrom, André Hinkenjann, Miguel L. Encarnação, Cláudio T. Silva, and Daniel Coming, editors. Advances in Visual Computing: 5th International Symposium, ISVC 2009, Las Vegas, NV, USA, November 30 — December 2, 2009, Proceedings, Part I, volume 5875 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10330-8 (print), 3-642-10331-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        10331-5.

Bebis:2010:AVCa

[BBP^{+}10a] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Ronald Chung, Riad Hammoud, Muhammad Hussain, Tan Kar-Han, Roger Crawfis, Daniel Thalmann, David Kao, and Lisa Avila, editors. Advances in Visual Computing: 6th International Symposium, ISVC 2010, Las Vegas, NV, USA, November 29–December 1, 2010. Proceedings, Part I, volume 6453 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17288-1 (print), 3-642-17289-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17289-2.

Bebis:2010:AVCb

[BBP^{+}10b] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Ronald Chung, Riad Hammound, Muhammad Hussain, Tan Kar-Han, Roger Crawfis, Daniel Thalmann, David Kao, and Lisa Avila, editors.

Advances in Visual Computing: 6th International Symposium, ISVC 2010, Las Vegas, NV, USA, November 29 — December 1, 2010, Proceedings, Part II, volume 6454 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17273-3 (print), 3-642-17274-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17274-8.

Bebis:2010:AVCc

[BBP^{+}10c] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Ronald Chung, Riad Hammound, Muhammad Hussain, Tan Kar-Han, Roger Crawfis, Daniel Thalmann, David Kao, and Lisa Avila, editors. Advances in Visual Computing: 6th International Symposium, ISVC 2010, Las Vegas, NV, USA, November 29 — December 1, 2010, Proceedings, Part III, volume 6455 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17276-8 (print), 3-642-17277-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17277-9.

Bebis:2011:AVCa

[BBP^{+}11a] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Song Wang, Kim Kyungnam, Bedrich Benes, Kenneth Moreland, Christoph Borst, Stephen Di-Verdi, Chiang Yi-Jen, and Jiang Ming, editors. Advances in Visual Computing: 7th International Symposium, ISVC 2011, Las Vegas, NV, USA, September 26– 28, 2011. Proceedings, Part I, volume 6938 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24027-5 (print), 3-642-24028-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24028-7.

Bebis:2011:AVCb

[BBP^{+}11b] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Song Wang, Kim Kyungnam, Bedrich Benes, Kenneth Moreland, Christoph Borst, Stephen Di-Verdi, Chiang Yi-Jen, and Jiang Ming, editors. Advances in Visual Computing: 7th International Symposium, ISVC 2011, Las Vegas, NV, USA, September 26– 28, 2011. Proceedings, Part II, volume 6939 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24030-5 (print), 3-642-24031-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24031-7.

Bebis:2012:AVCa

[BBP^{+}12a] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Charless Fowlkes, Sen Wang, Min-Hyung Choi, Stephan Mantler, Jürgen Schulze, Daniel Acevedo, Klaus Mueller, and Michael Papka, editors. Advances in Visual Computing: 8th International Symposium, ISVC 2012, Rethymnon, Crete, Greece, July 16–18, 2012, Revised Selected Papers, Part I, volume 7431 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33178-5 (print), 3-642-33179-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33179-4.

Bebis:2012:AVCb

[BBP^{+}12b] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Charless Fowlkes, Sen Wang, Min-Hyung Choi, Stephan Mantler, Jürgen Schulze, Daniel Acevedo, Klaus Mueller, and Michael Papka, editors. Advances in Visual Computing: 8th International Symposium, ISVC 2012, Rethymnon, Crete, Greece, July 16–18, 2012, Revised Selected Papers, Part II, volume 7432 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33190-4 (print), 3-642-33191-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-642-
33191-6.

Bove:2009:LER

[BBPP09] Ana Bove, Luís Soares Barbosa, Alberto Pardo, and Jorge Sousa Pinto, editors. Language Engineering and Rigorous Software Development: International LerNet ALFA Summer School 2008, Piriapolis, Uruguay, February 24 — March 1, 2008, Revised Tutorial Lectures, volume 5520 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03152-8 (print), 3-642-03153-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03153-3.

Borger:1988:CWC

[BBR88] E. Borger, H. Kleine Buning, and Michael M. Richter, editors. CSL ’87: 1st Workshop on Computer Science Logic, Karlsruhe, FRG, October 12–16, 1987: proceedings, volume 329 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-50241-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W6241 1987. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0329.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        50241-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=329.

Borger:1989:CWC

[BBR89] E. Borger, H. Kleine Buning, and Michael M. Richter, editors. CSL ’88: 2nd Workshop on Computer Science Logic, Duisburg, FRG, October 3–7, 1988: proceedings, volume 385 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51659-X (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 W67 1988. DM57.50. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0385.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51659-2;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=385.

Borger:1990:CWC

[BBR90] E. Borger, H. Kleine Buning, and Michael M. Richter, editors. CSL ’89: 3rd Workshop on Computer Science Logic, Kaiserslautern, FRG, October 2–6, 1989: proceedings, volume 440 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52753-2, 3-540-52753-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 W67 1989. DM57.50. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0440.htm;
http://www.springerlink.
com/content/978-0-387-
52753-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=440.

Bondavalli:2007:DCT

[BBR07] Andrea Bondavalli, Francisco Brasileiro, and Sergio Rajsbaum, editors. Dependable Computing: Third Latin-American Symposium, LADC 2007, Morella, Mexico, September 26– 28, 2007. Proceedings, volume 4746 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75293-5 (print), 3-540-75294-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75294-3.

Blesa:2010:HMI

[BBR^{+}10] María J. Blesa, Christian Blum, Günther Raidl, Andrea Roli, and Michael Sampels, editors. Hybrid Metaheuristics: 7th International Workshop, HM 2010, Vienna, Austria, October 1–2, 2010. Proceedings, volume 6373 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16053-0 (print), 3-642-16054-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16054-7.

Bixby:1998:IPC

[BBRM98] Robert E. Bixby, E. Andrew Boyd, and Roger Z. Rios-Mercado, editors. Integer programming and combinatorial optimization: 6th International IPCO Conference, Houston, Texas, June 22– 24, 1998: Proceedings, volume 1412 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64590-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1412. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1412.htm;
 http://www.springerlink.
com/content/978-3-540-
64590-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1412.

Blesa:2005:HMS

[BBRS05] María J. Blesa, Christian Blum, Andrea Roli, and Michael Sampels, editors. Hybrid Metaheuristics: Second International Workshop, HM 2005, Barcelona, Spain, August 29–30, 2005. Proceedings, volume 3636 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28535-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28535-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3636.

Baral:2007:LPN

[BBS07] Chitta Baral, Gerhard Brewka, and John Schlipf, editors. Logic Programming and Nonmonotonic Reasoning: 9th International Conference, LP-NMR 2007, Tempe, AZ, USA, May 15–17, 2007. Proceedings, volume 4483 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72199-1 (print), 3-540-72200-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72200-7.

Baker:2008:PDS

[BBT08] Theodore P. Baker, Alain Bui, and Sébastien Tixeuil, editors. Principles of Distributed Systems: 12th International Conference, OPODIS 2008, Luxor, Egypt, December 15–18, 2008. Proceedings, volume 5401 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92220-2 (print), 3-540-92221-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92221-6.

Braubach:2010:PMA

[BBT10] Lars Braubach, Jean-Pierre Briot, and John Thangarajah, editors. Programming Multi-Agent Systems: 7th International Workshop, Pro-MAS 2009, Budapest, Hungary, May 10–15, 2009. Revised Selected Papers, volume 5919 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14842-5 (print), 3-642-14843-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14843-9.

Baldoni:2010:DAL

[BBvRL10] Matteo Baldoni, Jamal Bentahar, M. Birna van Riemsdijk, and John Lloyd, editors. Declarative Agent Languages and Technologies VII:

7th International Workshop, DALT 2009, Budapest, Hungary, May 11, 2009. Revised Selected and Invited Papers, volume 5948 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11354-0 (print), 3-642-11355-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11355-0.

Bjorner:2000:PSI

[BBZ00] D. Bjørner, M. Broy, and A. V. Zamulin, editors. Perspectives of system informatics: Third International Andrei Ershov Memorial Conference, PSI ’99, Akademgorodok, Novosibirsk, Russia, July 6–9, 1999: Proceedings, volume 1755 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67102-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .I497 2000; QA267.A1 L43 no.1755.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1755.

htm;   http://www.springerlink.
com/content/978-3-540-
67102-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1755.

Bjorner:2001:PSI

[BBZ01] Dines Bjørner, Manfred Broy, and Alexandre V. Zamulin, editors. Perspectives of system informatics: 4th International Andrei Ershov Memorial Conference, PSI ’2001, Akademgorodok, Novosibirsk, Russia, July 2– 6, 2001: Revised Papers, volume 2244 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43075-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .A454 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2244.htm;
http://www.springerlink.
com/content/978-3-540-
43075-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2244.

Blakley:1985:ACP

[BC85] George Robert Blakley and David Chaum, editors. Advances in Cryptology: Proceedings of CRYPTO 84, volume 196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CODEN LNCSD9. ISBN 0-387-15658-5, 3-540-39568-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C791 1984; QA267.A1 L43 no.196. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0196.htm;
 http://www.springerlink.
com/content/978-0-387-
15658-3;   http://www.
springerlink.com/content/
cemajg0qmeev/;   http:
//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=196. CRYPTO

84: a Workshop on the Theory and Application of Cryptographic Techniques, held at the University of California, Santa Barbara, August 19–22, 1984, sponsored by the International Association for Cryptologic Research.

Beth:1988:AAE

[BC88] Thomas Beth and Michael Clausen, editors. Applicable algebra, errorcorrecting codes, combinatorics and computer algebra: proceedings, volume 307 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19200-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A351 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0307.htm;
 http://www.springerlink.
com/content/978-0-387-
19200-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=307.

Bagchi:1989:IRD

[BC89] Tapan P. Bagchi and Vinay K. Chaudhri. Interactive relational database design: a logic programming implementation, volume 402 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51881-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 186 pp. LCCN QA76.9.D26 B34 1989.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0402.
htm;   http://www.springerlink.
com/content/978-0-387-
51881-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=402.

Bekkers:1992:MMI

[BC92] Y. Bekkers and J. Cohen, editors. Memory management: International Workshop IWMM ’92, St. Malo, France, September 17–19, 1992: proceedings, volume 637 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55940-X (Berlin), 0-387-55940-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M45 I96 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0637.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55940-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=637.

Errata slip inserted.

Bidoit:1993:RTD

[BC93] M. Bidoit and C. Choppy, editors. Recent trends in data type specification: 8th Workshop on Specification of Abstract Data Types, joint with the 3rd COMPASS Workshop, Dourdan, France, August 26–30, 1991: Selected Papers, volume 655 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56379-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A23 W67 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0655.
htm;   http://www.springerlink.
com/content/978-0-387-
56379-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=655.

Balas:1995:IPC

[BC95] Egon Balas and Jens Clausen, editors. Integer programming and combinatorial optimization: 4th International ICPO Conference, Copenhagen, Denmark, May 29–31, 1995: proceedings, volume 920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59408-6 (Berlin), 0-387-59408-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T57.74 .C6 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0920.
htm;   http://www.springerlink.
com/content/978-3-540-
59408-6;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=920.

Berardi:1996:TPP

[BC96a] Stefano Berardi and Mario Coppo, editors. Types for proofs and programs: International Workshop TYPES ’95, Torino, Italy, June 5– 8, 1995: Selected Papers, volume 1158 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61780-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T96 1995.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1158.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61780-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1158.

Buxton:1996:CVEa

[BC96b] Bernard F. Buxton and Roberto Cipolla, editors. Computer vision, ECCV ’96: Fourth European Conference on Computer Vision, Cambridge, UK, April 15– 18, 1996: Proceedings. Part I, volume 1064 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61122-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E97 1996 v.1-2 (c1996). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1064.htm;
 http://www.springerlink.
com/content/978-3-540-
61122-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1064.

Buxton:1996:CVEb

[BC96c] Bernard F. Buxton and Roberto Cipolla, editors. Computer vision, ECCV ’96: Fourth European Conference on Computer Vision, Cambridge, UK, April 15– 18, 1996: Proceedings. Part II, volume 1065 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61123-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E97 1996 v.1-2 (c1996). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1065.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        61123-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1065.

Burke:1998:PTA

[BC98] Edmund Burke and Michael Carter, editors. Practice and theory of automated timetabling II: Second International Conference, PATAT’97, Toronto, Canada, August 1997: Selected Papers, volume 1408 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64979-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q340.I58 1997.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1408.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        64979-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1408.

Borner:2002:VID

[BC02] Katy Börner and Chaomei Chen, editors. Visual Interfaces to Digital Libraries, volume 2539 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00247-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.U83 V57 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2539.htm;
 http://www.springerlink.
com/content/978-3-540-
00247-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2539.

Papers presented at the first and Second International Workshops on “Visual Interfaces to Digital Libraries” held at the Joint Conference on Digital Libraries (JCDL) in 2001 and 2002.

Bernardo:2004:FMD

[BC04a] Marco Bernardo and Flavio Corradini, editors. Formal methods for the design of real-time systems: international school on formal methods for the design of computer, communication, and software systems, Bertinora, Italy, September 13–18, 2004, revised lectures, volume 3185 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23068-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.54 .I575 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        23068-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3185;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b110123.

Buchberger:2004:AIS

[BC04b] Bruno Buchberger and John A. Campbell, editors. Artificial intelligence and symbolic computation: 7th international conference, AISC 2004, Linz, Austria, September 22–24, 2004: proceedings, volume 3249 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23212-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A385 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        23212-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3249;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100361.

Blundo:2005:SCN

[BC05] Carlo Blundo and Stelvio Cimato, editors. Security in communication networks: 4th international conference, SCN 2004, Amalfi, Italy, September 8–10, 2004: Revised selected papers, volume 3352 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24301-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.59 .S385 2004.

URL http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3352;
http://www.springerlink.
com/content/978-3-540-
24301-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3352;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105083.

Bersini:2006:AIS

[BC06] Hugues Bersini and Jorge Carneiro, editors. Artificial Immune Systems: 5th International Conference, ICARIS 2006, Oeiras, Portugal, September 4–6, 2006. Proceedings, volume 4163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37749-2 (print), 3-540-37751-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37751-1.

Bonato:2007:AMW

[BC07] Anthony Bonato and Fan R. K. Chung, editors. Algorithms and models for the web-graph: 5th international workshop, WAW 2007, San Diego, CA, USA, December 11–12, 2007: proceedings, volume 4863 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-77003-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W39 2007; QA76.9.A43 W428 2007; QA76.9.A43 W428 2007; Internet; QA76.9.A43 W63 2007. URL http:

//www.springerlink.
com/content/978-3-540-
77003-9.

Borger:2008:ASE

[BC08] Egon Börger and Antonio Cisternino, editors. Advances in Software Engineering: Lipari Summer School 2007, Lipari Island, Italy, July 8–21, 2007, Revised Tutorial Lectures, volume 5316 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89761-5 (print), 3-540-89762-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89762-0.

Bacon:2009:MAI

[BC09a] Jean M. Bacon and Brian F. Cooper, editors. Middleware 2009: ACM/ IFIP/USENIX, 10th International Middleware Conference, Urbana, IL, USA, November 30 — December 4, 2009. Proceedings, volume 5896 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10444-4 (print), 3-642-10445-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10445-9.

Breitman:2009:FMS

[BC09b] Karin Breitman and Ana Cavalcanti, editors. Formal Methods and Software Engineering: 11th International Conference on Formal Engineering Methods ICFEM 2009, Rio de Janeiro, Brazil, December 9–12, 2009. Proceedings, volume 5885 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10372-3 (print), 3-642-10373-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10373-5.

Bello:2010:BSI

[BC10a] Fernando Bello and Stéphane Cotin, editors. Biomedical Simulation: 5th International Symposium, ISBMS 2010, Phoenix, AZ, USA, January 23–24, 2010. Proceedings, volume 5958 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11614-0 (print), 3-642-11615-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11615-5.

Bloch:2010:PPR

[BC10b] Isabelle Bloch and Roberto M.

Cesar Jr., editors. Progress in Pattern Recognition, Image Analysis, Computer Vision, and Applications: 15th Iberoamerican Congress on Pattern Recognition, CIARP 2010, Sao Paulo, Brazil, November 8–11, 2010. Proceedings, volume 6419 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16686-5 (print), 3-642-16687-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16687-7.

Bouajjani:2010:ATV

[BC10c] Ahmed Bouajjani and Wei-Ngan Chin, editors. Automated Technology for Verification and Analysis: 8th International Symposium, ATVA 2010, Singapore, September 21–24, 2010. Proceedings, volume 6252 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15642-8 (print), 3-642-15643-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15643-4.

Bernstein:2011:PCI

[BC11a] Daniel J. Bernstein and Sanjit Chatterjee, editors. Progress in Cryptology — INDOCRYPT 2011: 12th International Conference on Cryptology in India, Chennai, India, December 11–14. Proceedings, volume 7107 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25577-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25577-9.

Boyen:2011:PSI

[BC11b] Xavier Boyen and Xiaofeng Chen, editors. Provable Security: 5th International Conference, ProvSec 2011, Xi’an, China, October 16–18, 2011. Proceedings, volume 6980 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24315-0 (print), 3-642-24316-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24316-5.

Brandes:2011:GDI

[BC11c] Ulrik Brandes and Sabine Cornelsen, editors. Graph Drawing: 18th International Symposium, GD 2010, Konstanz, Germany, September 21–24, 2010. Revised Selected Papers, volume 6502 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18468-5 (print), 3-642-18469-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18469-7.

Bigun:1997:AVB

[BCB97] Josef Bigun, Gerard Chollet, and Gunilla. Borgefors, editors. Audioand video-based biometric person authentication: First International Conference, AVBPA ’97, Crans-Montana, Switzerland, March 12–14, 1997: proceedings, volume 1206 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62660-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.S65 A944 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1206.htm;
 http://www.springerlink.
com/content/978-3-540-
62660-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1206.

Bal:1999:IPL

[BCB99] Henri Bal, Luca Cardelli, and Boumediene Belkhouche, editors. Internet programming languages: ICCL’98 Workshop, held in Chicago, IL, USA, May 13, 1998: Proceedings, volume 1686 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66673-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.625 I23 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1686.htm;
 http://www.springerlink.
com/content/978-3-540-

66673-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1686.

Bairoch:2008:DIL

[BCBF08] Amos Bairoch, Sarah Cohen-Boulakia, and Christine Froidevaux, editors. Data Integration in the Life Sciences: 5th International Workshop, DILS 2008, Evry, France, June 25– 27, 2008. Proceedings, volume 5109 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69827-2 (print), 3-540-69828-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69828-9.

Benjamins:2005:LSW

[BCBG05] V. Richard Benjamins, Pompeu Casanovas, Joost Breuker, and Aldo Gangemi, editors. Law and the Semantic Web: legal ontologies, methodologies, legal information retrieval, and applications, volume 3369 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25063-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815 .L39 2005. URL http:

//www.springerlink.
com/content/978-3-540-
25063-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3369;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106624.

Burgard:1999:KAA

[BCC99] Wolfram Burgard, Thomas Christaller, and A. B. Cremers, editors. KI-99: advances in artificial intelligence: 23rd Annual German Conference on Artificial Intelligence, Bonn, Germany, September 13–15, 1999: Proceedings, volume 1701 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66495-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1701; Q334 .G47 1999. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t1701.htm;
http://www.springerlink.
com/content/978-3-540-
66495-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1701.

Barkaoui:2006:TAC

[BCC06] Kamel Barkaoui, Ana Cavalcanti, and Antonio Cerone, editors. Theoretical Aspects of Computing — ICTAC 2006: Third International Colloquium, Tunis, Tunisia, November 20–24, 2006. Proceedings, volume 4281 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48815-4 (print), 3-540-48816-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48816-3.

Borrajo:2007:CTA

[BCC07] Daniel Borrajo, Luis Castillo, and Juan Manuel Corchado, editors. Current Topics in Artificial Intelligence: 12th Conference of the Spanish Association for Artificial Intelligence, CAEPIA 2007, Salamanca, Spain, November 12–16, 2007.

Selected Papers, volume 4788 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75270-6 (print), 3-540-75271-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75271-4.

Bellavista:2010:AGP

[BCC^{+}10] Paolo Bellavista, Ruay-Shiung Chang, Han-Chieh Chao, Shin-Feng Lin, and Peter M. A. Sloot, editors. Advances in Grid and Pervasive Computing: 5th International Conference, GPC 2010, Hualien, Taiwan, May 10–13, 2010. Proceedings, volume 6104 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13066-6 (print), 3-642-13067-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13067-0.

Boissonnat:2012:CSI

[BCC^{+}12] Jean-Daniel Boissonnat, Patrick Chenin, Albert Cohen, Christian Gout, Tom Lyche, Marie-Laurence Mazure, and Larry Schumaker, editors. Curves and Surfaces: 7th International Conference, Avignon, France, June 24 — 30, 2010, Revised Selected Papers, volume 6920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27412-9 (print), 3-642-27413-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27413-8.

Boyd:2000:ISP

[BCD00] Colin Boyd, Andrew Clark, and Ed Dawson, editors. Information security and privacy: 5th Australasian Conference, ACISP 2000, Brisbane, Australia, July 10–12, 2000: Proceedings, volume 1841 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67742-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1841.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1841.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67742-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1841.

Banzhaf:2003:AAL

[BCD^{+}03] Wolfgang Banzhaf, Thomas Christaller, Peter Dittrich, Jan T. Kim, and Jens Ziegler, editors. Advances in Artificial Life: 7th European Conference, ECAL 2003 Dortmund, Germany, September 14–17, 2003 Proceedings, volume 2801 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20057-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .E87 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2801.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20057-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2801;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12035.

Berardi:2004:TPP

[BCD04] Stefano Berardi, Mario Coppo, and Ferruccio Damiani, editors. Types for Proofs and Programs: International Workshop, Types 2003, Torino, Italy, April 30–May 4, 2003, Revised Selected Papers, volume 3085 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22164-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I557 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3085.htm;
http://www.springerlink.
com/content/978-3-540-
22164-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3085;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98246.

Bento:2005:PAI

[BCD05] Carlos Bento, Amílcar Cardoso, and Gaël Dias, editors. Progress in artificial intelligence: 12th Portuguese Conference on Artificial Intelligence, EPIA 2005, Covilhã, Portugal, December 5– 8, 2005: proceedings, volume 3808 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30737-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30737-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3808.

Bussler:2007:BIR

[BCDN07] Christoph Bussler, Malu Castellanos, Umesh Dayal, and Sham Navathe, editors. Business Intelligence for the Real-Time Enterprises: First International Workshop, BIRTE 2006, Seoul, Korea, September 11, 2006, Revised Selected Papers, volume 4365 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73949-1 (print), 3-540-73950-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73950-0.

Bayro-Corrochano:2009:PPR

[BCE09] Eduardo Bayro-Corrochano and Jan-Olof Eklundh, editors. Progress in Pattern Recognition, Image Analysis, Computer Vision, and Applications: 14th Iberoamerican Conference on Pattern Recognition, CIARP 2009, Guadalajara, Jalisco, Mexico, November 15– 18, 2009. Proceedings, volume 5856 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10267-0 (print), 3-642-10268-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10268-4.

Buchmann:2002:TST

[BCF^{+}02] Alejandro P. Buchmann, Fabio Casati, Ludger Fiege, Mei-Chun Hsu, and Ming-Chien Shan, editors. Technologies for E-Services: Third International Workshop, TES 2002, Hong Kong, China, August 23–24, 2002. Proceedings, volume 2444 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44110-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .T43 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2444.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44110-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2444.

Braun:2006:WWI

[BCF06] Torsten Braun, Georg Carle, and Sonia Fahmy, editors. Wired/Wireless Internet Communications: 4th International Conference, WWIC 2006, Bern, Switzerland, May 10–12, 2006, Proceedings, volume 3970 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34023-8 (softcover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34023-2.

Backhouse:2002:ACM

[BCG02] Roland C. Backhouse, Roy L. Crole, and Jeremy Gibbons, editors. Algebraic and coalgebraic methods in the mathematics of program construction: international summer school and workshop Oxford, UK, April 2000, revised lectures, volume 2297 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43613-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S415 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2297.htm;
 http://www.springerlink.
com/content/978-3-540-
43613-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2297.

Benatallah:2007:WIS

[BCG^{+}07]  Boualem  Benatallah,

Fabio Casati, Dimitrios Georgakopoulos, Claudio Bartolini, Wasim Sadiq, and Claude Godart, editors. Web Information Systems Engineering — WISE 2007: 8th International Conference on Web Information Systems Engineering Nancy, France, December 3–7, 2007 Proceedings, volume 4831 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76992-7 (print), 3-540-76993-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76993-4.

Bernardi:2011:ALT

[BCG^{+}11] Raffaella Bernardi, Sally Chambers, Björn Gottfried, Frédérique Segond, and Ilya Zaihrayeu, editors. Advanced Language Technologies for Digital Libraries: International Workshops on NLP4DL 2009, Viareggio, Italy, June 15, 2009 and AT4DL 2009, Trento, Italy, September 8, 2009, volume 6699 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23159-4 (print), 3-642-23160-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23160-5.

Borgida:2009:CMF

[BCGY09] Alexander T. Borgida, Vinay K. Chaudhri, Paolo Giorgini, and Eric S. Yu, editors. Conceptual Modeling: Foundations and Applications: Essays in Honor of John Mylopoulos, volume 5600 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02462-9 (print), 3-642-02463-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02463-4.

Bressan:2005:DXT

[BCH^{+}05] Stéphane Bressan, Stefano Ceri, Ela Hunt, Zachary G. Ives, Zohra Bellahsène, Michael Rys, and Rainer Unland, editors. Database and XML technologies: Third International XML Database Symposium, XSym 2005, Trondheim, Norway, August 28–29, 2005: proceedings, volume 3671 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28583-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3.I5978 2005. URL http:

//www.springerlink.
com/content/978-3-540-
28583-0;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3671.

Beth:1985:ACP

[BCI85] Thomas Beth, N. Cot, and I. Ingemarsson, editors. Advances in cryptology: proceedings of EUROCRYPT 84, a Workshop on the Theory and Application of Cryptographic Techniques, Paris, France, April 9– 11, 1984, volume 209 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-16076-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E951 1984. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0209.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        16076-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=209.

Held at the University of

Paris, Sorbonne.

Bennedsen:2008:RTP

[BCK08] Jens Bennedsen, Michael E. Caspersen, and Michael Kölling, editors. Reflections on the Teaching of Programming: Methods and Implementations, volume 4821 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77933-7 (print), 3-540-77934-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77934-6.

Boulic:2010:MGT

[BCK10] Ronan Boulic, Yiorgos Chrysanthou, and Taku Komura, editors. Motion in Games: Third International Conference, MIG 2010, Utrecht, The Netherlands, November 14–16, 2010. Proceedings, volume 6459 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16957-0 (print), 3-642-16958-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16958-8.

Barni:2005:DWI

[BCKK05] Mauro Barni, Ingemar Cox, Ton Kalker, and Hyoung Joong Kim, editors. Digital watermarking: 4th international workshop, IWDW 2005, Siena, Italy, September 15–17, 2005: proceedings, volume 3710 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28768-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I939 2005.

URL http://www.springerlink.
com/content/978-3-540-
28768-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3710.

Benatallah:2010:WEI

[BCKR10]  Boualem  Benatallah,

Fabio Casati, Gerti Kappel, and Gustavo Rossi, editors. Web Engineering: 10th International Conference, ICWE 2010, Vienna Austria, July 5–9, 2010. Proceedings, volume 6189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13910-8 (print), 3-642-13911-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13911-6.

Braun:2005:WWI

[BCKT05] Torsten Braun, Georg Carle, Yevgeni Koucheryavy, and Vassilis Tsaoussidis, editors. Wired/ wireless Internet communications: third international conference, WWIC 2005, Xanthi, Greece, May 11–13, 2005. Proceedings, volume 3510 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25899-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25899-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3510;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136247.

Battiti:2003:WDN

[BCL03] Roberto Battiti, Marco Conti, and Renato Lo Cigno, editors. Wireless On-Demand Network Systems: First IFIP TC6 Working Conference, WONS 2004, Madonna di Campiglio, Italy, January 21–23, 2004: Proceedings, volume 2928 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20790-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.78. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2928.htm;
 http://www.springerlink.
com/content/978-3-540-
20790-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2928;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94064.

Bert:2000:RTA

[BCM00] Didier Bert, C. Choppy, and Peter Mosses, editors. Recent trends in algebraic development techniques: 14th International Workshop, WADT’99, Chateau de Bonas, September 15– 18, 1999: Selected Papers, volume 1827 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67898-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1827. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1827.htm;
http://www.springerlink.
com/content/978-3-540-
67898-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1827.

Bazzan:2008:ABC

[BCM08] Ana L. C. Bazzan, Mark Craven, and Natália F. Martins, editors. Advances in Bioinformatics and Computational Biology: Third Brazilian Symposium on Bioinformatics, BSB 2008, Santo André, Brazil, August 28–30, 2008. Proceedings, volume 5167 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85556-4 (print), 3-540-85557-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85557-6.

Bonuccelli:1994:ACS

[BCP94] M. Bonuccelli, Pierluigi Crescenzi, and Rosella Petreschi, editors. Algorithms and complexity: second Italian conference, CIAC ’94, Rome, Italy, February 23–25, 1994: proceedings, volume 778 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57811-0 (New York), 3-540-57811-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .A42 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0778.htm;
 http://www.springerlink.
com/content/978-0-387-
57811-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=778.

Papers presented at the Second Italian Conference on Algorithms and Complexity (CIAC ’94).

Bertrand:1999:DGC

[BCP99] Gilles Bertrand, Michel Couprie, and Laurent Perroton, editors. Discrete geometry for computer imagery: 8th International Conference, DGCI ’99, Marne-la-Vallée, France, March 17–19, 1999: Proceedings, volume 1568 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65685-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .D525 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1568.htm;
 http://www.springerlink.
com/content/978-3-540-
65685-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1568.

Bertino:2004:ADT

[BCP^{+}04] Elisa Bertino, Stavros Christodoulakis, Dimitris Plexousakis, Vassilis Christophides, Manolis Koubarakis, Klemens Böhm, and Elena Ferrari, editors. Advances in Database Technology— EDBT 2004: 9th International Conference on Extending Database Technology Heraklion, Crete, Greece, March 14–18, 2004 Proceedings, volume 2992 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21200-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5534 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2992.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        21200-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2992;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95855.

Bernardo:2012:FMM

[BCP12] Marco Bernardo, Vittorio Cortellessa, and Alfonso Pierantonio, editors. Formal Methods for Model-Driven Engineering: 12th International School on Formal Methods for the Design of Computer, Communication, and Software Systems, SFM 2012, Bertinoro, Italy, June 18– 23, 2012. Advanced Lectures, volume 7320 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30981-X (print), 3-642-30982-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30982-3.

Brusilovsky:2003:UMI

[BCR03] Peter Brusilovsky, Albert T. Corbett, and Fiorella De Rosis, editors. User Modeling 2003: 9th International Conference, UM 2003, Johnstown, PA, USA, June 22–26, 2003: Proceedings, volume 2702 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40381-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I585 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2702.htm;
 http://www.springerlink.
com/content/978-3-540-
40381-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2702.

Beigl:2011:MUC

[BCRB^{+}11] Michael Beigl, Henning Christiansen, Thomas R. Roth-Berghofer, Anders Kofod-Petersen, Kenny R. Coventry, and Hedda R. Schmidtke, editors. Modeling and Using Context: 7th International and Interdisciplinary Conference, CONTEXT 2011, Karlsruhe, Germany, September 26–30, 2011. Proceedings, volume 6967 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24278-2 (print), 3-642-24279-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24279-3.

Bouge:1992:PPC

[BCRT92] L. Bouge, M. Cosnard, Y. Robert, and D. Trystram, editors. Parallel processing: CON-PAR 92-VAPP V, Second Joint International Conference on Vector and Parallel Processing, Lyon, France, September 1–4, 1992: proceedings, volume 634 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55895-0 (Berlin), 0-387-55895-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .J65 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0634.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55895-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=634.

Baresi:2009:SOC

[BCS09] Luciano Baresi, Chi-Hung Chi, and Jun Suzuki, editors. Service-Oriented Computing: 7th International Joint Conference, ICSOC-ServiceWave 2009, Stockholm, Sweden, November 24–27, 2009. Proceedings, volume 5900 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10382-0 (print), 3-642-10383-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10383-4.

Bench-Capon:1999:DES

[BCST99] T. J. M. (Trevor J. M.)

Bench-Capon, Giovanni Soda, and A. Min Tjoa, editors. Database and expert systems applications: 10th International Conference, DEXA’99, Florence, Italy, August 30–September 3, 1999: Proceedings, volume 1677 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66448-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D29 1999.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1677.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66448-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1677.

Borba:2010:TTS

[BCSW10] Paulo Borba, Ana Cavalcanti, Augusto Sampaio, and Jim Woodcook, editors. Testing Techniques in Software Engineering: Second Pernambuco Summer School on Software Engineering, PSSE 2007, Recife, Brazil, December 3– 7, 2007, Revised Lectures, volume 6153 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14334-2 (print), 3-642-14335-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14335-9.

Bandini:2002:CAI

[BCT02a] S. Bandini, Bastien Chopard, and Marco Tomassini, editors. Cellular Automata: 5th International Conference on Cellular Automata for Research and Industry, ACRI 2002, Geneva, Switzerland, October 9– 11, 2002. Proceeding, volume 2493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44304-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.C45 I58 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2493.
htm;   http://www.springerlink.
com/content/978-3-540-
44304-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2493.

Also available via the World Wide Web.

Batory:2002:GPC

[BCT02b] Don S. Batory, C. (Charles)

Consel, and Walid Taha, editors. Generative Programming and Component Engineering: ACM SIGPLAN/SIGSOFT Conference, GPCE 2002, Pittsburgh, PA, USA, October 6–8, 2002. Proceedings, volume 2487 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44284-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.624 .G63 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2487.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44284-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2487.

Birk:2002:RRS

[BCT02c] Andreas Birk, Silvia Coradeschi, and Satoshi Tadokoro, editors. RoboCup 2001: Robot Soccer World Cup V, volume 2377 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43912-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ211 .R532 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2377.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43912-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2377.

Benatallah:2006:SOC

[BCT06]  Boualem  Benatallah,

Fabio Casati, and Paolo Traverso, editors. Serviceoriented computing — ICSOC 2005: third international conference, Amsterdam, The Netherlands, December 12– 15, 2005, proceedings, volume 3826 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30817-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30817-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3826.

Bertels:2006:RCA

[BCV06] Koen Bertels, João M. P.

Cardoso, and Stamatis Vassiliadis, editors. Reconfigurable Computing: Architectures and Applications: Second International Workshop, ARC 2006, Delft, The Netherlands, March 1– 3, 2006, Revised Selected Papers, volume 3985 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36708-X (print), 3-540-36863-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        36863-2.

Bode:1993:PCA

[BD93] Arndt Bode and Mario Dal Cin. Parallel computer architectures: theory, hardware, software, applications, volume 732 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57307-0 (Berlin), 0-387-57307-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 311 pp. LCCN QA76.9.A73 P39 1993. DM66.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0732.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        57307-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=732.

Ben-David:1997:CLT

[BD97a] Shai Ben-David, editor. Computational learning theory: Third European Conference, Euro-COLT’97, Jerusalem, Israel, March 17–19, 1997 proceedings, volume 1208 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62685-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.7 .E97 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1208.htm;
 http://www.springerlink.
com/content/978-3-540-
62685-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1208.

Bidoit:1997:TTP

[BD97b] M. Bidoit and M. Dauchet, editors. TAPSOFT ’97: theory and practice of software development: 7th International Joint Conference CAAP/FASE, Lille, France, April 14–18, 1997: proceedings, volume 1214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62781-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1214.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1214.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        62781-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1214.

Birk:1998:LRE

[BD98] Andreas Birk and John Demiris, editors. Learning robots: 6th European Workshop, EWLR-6, Brighton, England, August 1–2, 1997: Proceedings, volume 1545 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65480-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58A97 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1545.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65480-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1545.

Burke:2003:PTA

[BD03] Edmund Burke and Patrick De Causmaecker, editors. Practice and Theory of Automated Timetabling IV: 4th International Conference, PATAT 2002, Gent, Belgium, August 21–23, 2002: revised Selected Papers, volume 2740 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40699-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q340 .P75 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2740.htm;
 http://www.springerlink.
com/content/978-3-540-
40699-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2740;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11828.

Brewster:2004:MHC

[BD04] Stephen Brewster and Mark Dunlop, editors. Mobile Human-Computer Interaction—MobileHCI 2004: 6th International Symposium, Glasgow, UK, September 13–16, 2004. Proceedings, volume 3160 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23086-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3160.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23086-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3160;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100594.

Buchmann:2008:PQC

[BD08] Johannes Buchmann and Jintai Ding, editors. Post-Quantum Cryptography: Second International Workshop, PQCrypto 2008 Cincinnati, OH, USA, October 17–19, 2008 Proceedings, volume 5299 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. ISBN 3-540-88403-3. LCCN QA76.9.A25 P63 2008.

                                        URL http://www.springer.
                                        com/computer/security+
                                        and+cryptology/book/
                                        978-3-540-88402-6;
                                        http://www.springerlink.
                                        com/content/978-3-540-
                                        88403-3.

Bruni:2011:FTD

[BD11] Roberto Bruni and Juergen Dingel, editors. Formal Techniques for Distributed Systems: Joint 13th IFIP WG 6.1 International Conference, FMOODS 2011, and 30th IFIP WG 6.1 International Conference, FORTE 2011, Reykjavik, Iceland, June 6–9, 2011. Proceedings, volume 6722 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21460-6 (print), 3-642-21461-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21461-5.

Bechet:2012:LAC

[BD12] Denis Béchet and Alexander Dikovsky, editors. Logical Aspects of Computational Linguistics: 7th International Conference, LACL 2012, Nantes, France, July 2– 4, 2012. Proceedings, volume 7351 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31261-6 (print), 3-642-31262-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31262-5.

Barthe:2008:FMO

[BdB08] Gilles Barthe and Frank S. de Boer, editors. Formal Methods for Open Object-Based Distributed Systems: 10th IFIP WG 6.1 International Conference, FMOODS 2008, Oslo, Norway, June 4– 6, 2008 Proceedings, volume 5051 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68862-5 (print), 3-540-68863-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68863-1.

Balsters:2001:DSE

[BdBC01] Herman Balsters, Bert de Brock, and S. Conrad, editors. Database schema evolution and meta-modeling: 9th International Workshop on Foundations of Models and Languages for Data and Objects, FoMLaDO/ DEMM-2000, Dagstuhl Castle, Germany, September 18–21, 2000: Selected Papers, volume 2065 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42272-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5854 2000; QA267.A1 L43 no.2065. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2065.htm;
 http://www.springerlink.
com/content/978-3-540-
42272-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2065.

Bilgic:2003:FSS

[BdBK03] Taner Bilgic, Bernard de Baets, and Okyay Kaynak, editors. Fuzzy sets and systems, IFSA 2003: 10th International Fuzzy Systems Association World Congress, Istanbul, Turkey, June 30–July 2, 2003: Proceedings, volume 2715 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40383-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA248 .I37 2003.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t2715.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40383-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2715.

Blyth:2012:FCD

[BDC12] Jim Blyth, Sven Dietrich, and L. Jean Camp, editors. Financial Cryptography and Data Security: FC 2012 Workshops, USEC and WECSR 2012, Kralendijk, Bonaire, March 2, 2012, Revised Selected Papers, volume 7398 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34637-5 (print), 3-642-34638-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34638-5.

Ben-David:2004:ALT

[BDCM04] Shai Ben-David, John Case, and Akira Maruoka, editors. Algorithmic Learning Theory: 15th International Conference, ALT 2004, Padova, Italy, October 2–5, 2004. Proceedings, volume 3244 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23356-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3244.htm;
 http://www.springerlink.
com/content/978-3-540-
23356-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3244;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100989.

Berardi:2009:TPP

[BDd09] Stefano Berardi, Ferruccio Damiani, and Ugo de’Liguoro, editors. Types for Proofs and Programs: International Conference, TYPES 2008 Torino, Italy, March 26–29, 2008 Revised Selected Papers, volume 5497 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02443-2 (print), 3-642-02444-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        02444-3.

Bononi:2012:DCN

[BDDM12] Luciano Bononi, Ajoy K. Datta, Stéphane Devismes, and Archan Misra, editors. Distributed Computing and Networking: 13th International Conference, ICDCN 2012, Hong Kong, China, January 3–6, 2012. Proceedings, volume 7129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25958-8 (print), 3-642-25959-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25959-3.

Bordini:2005:PMA

[BDDS05] Rafael H. Bordini, Mehdi Dastani, Jürgen Dix, and Amal El Fallah Seghrouchni, editors. Programming Multi-Agent Systems: Second International Workshop PROMAS 2004, New York, NY, July 20, 2004, Selected Revised and Invited Papers, volume 3346 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24559-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 P79 2004. URL http:

//www.springerlink.
com/content/978-3-540-
24559-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3346;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105887.

Bordini:2006:PMA

[BDDS06] Rafael H. Bordini, Mehdi Dastani, Jürgen Dix, and Amal ElFallah Seghrouchni, editors. Programming Multi-Agent Systems: Third International Workshop, ProMAS 2005, Utrecht, The Netherlands, July 26, 2005, Revised and Invited Papers, volume 3862 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32616-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-

                                        32616-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3862.

Bordini:2007:PMA

[BDDS07] Rafael H. Bordini, Mehdi Dastani, Jürgen Dix, and Amal El Fallah Seghrouchni, editors. Programming Multi-Agent Systems: 4th International Workshop, Pro-MAS 2006, Hakodate, Japan, May 9, 2006, Revised and Invited Papers, volume 4411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71955-5 (print), 3-540-71956-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71956-4.

Barthe:2012:FAS

[BDE12] Gilles Barthe, Anupam Datta, and Sandro Etalle, editors. Formal Aspects of Security and Trust: 8th International Workshop, FAST 2011, Leuven, Belgium, September 12–14, 2011. Revised Selected Papers, volume 7140 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29419-7 (print), 3-642-29420-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29420-4.

Bodlaender:2012:MAR

[BDFM12] Hans L. Bodlaender, Rod Downey, Fedor V. Fomin, and Dániel Marx, editors. The Multivariate Algorithmic Revolution and Beyond: Essays Dedicated to Michael R. Fellows on the Occasion of His 60th Birthday, volume 7370 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30890-2 (print), 3-642-30891-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30891-8.

Bussler:2004:SWR

[BDFS04] Christoph Bussler, John Davies, Dieter Fensel, and Rudi Studer, editors. The Semantic Web: Research and Applications: First European Semantic Web Symposium, ESWS 2004, Heraklion, Crete, Greece, May 10–12, 2004: Proceedings, volume 3053 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21999-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815 .E97 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3053.
htm;   http://www.springerlink.
com/content/978-3-540-
21999-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3053;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97867.

Benczur:2004:ADI

[BDG04] András Benczúr, János Demetrovics, and Georg Gottlob, editors. Advances in Databases and Information Systems: 8th East European Conference, AD-BIS 2004, Budapest, Hungary, September 22– 25, 2004: Proceedings, volume 3255 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23243-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9 D3 A675 2004; QA76 A1 L43 3255. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3255.htm;
 http://www.springerlink.
com/content/978-3-540-
23243-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3255;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100480.

Beckert:2012:FVO

[BDG12a] Bernhard Beckert, Ferruccio Damiani, and Dilian Gurov, editors. Formal Verification of Object-Oriented Software: International Conference, FoVeOOS 2011, Turin, Italy, October 5– 7, 2011, Revised Selected Papers, volume 7421 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31761-8 (print), 3-642-31762-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31762-0.

Beneventano:2012:APP

[BDG^{+}12b] Domenico Beneventano, Zoran Despotovic, Francesco Guerra, Sam Joseph, Gianluca Moro, and Adrián Perreau de Pinninck, editors. Agents and Peer-to-Peer Computing: 7th International Workshop, AP2PC 2008, Estoril, Portugal, May 13, 2008 and 8th International Workshop, AP2PC 2009, Budapest, Hungary, May 11, 2009. Revised Selected Papers, volume 6573 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31808-8 (print), 3-642-31809-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31809-2.

Bowen:2000:ZFS

[BDGK00] J. P. Bowen, S. Dunne, A. Galloway, and S. King, editors. ZB 2000: Formal Specification and Development in Z and B: First International Conference of B and Z Users, York, UK, August 29– September 2, 2000. Proceedings, volume 1878 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67944-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.B155 I58 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1878.
htm;   http://www.springerlink.
com/content/978-3-540-
67944-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1878.

Baresi:2005:UMI

[BDGM05] Luciano Baresi, Schahram Dustdar, Harald Gall, and Maristella Matera, editors. Ubiquitous mobile information and collaboration systems: second CAISE workshop, UMICS 2004, Riga, Latvia, June 7– 8, 2004, revised selected papers, volume 3272 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24100-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
24100-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3272.

Boyanov:2007:NMA

[BDGN07] Todor Boyanov, Stefka Dimova, Krassimir Georgiev, and Geno Nikolov, editors. Numerical Methods and Applications: 6th International Conference, NMA 2006, Borovets, Bulgaria, August 20– 24, 2006. Revised Papers, volume 4310 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70940-1 (print), 3-540-70942-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70942-8.

Bertot:1999:TPH

[BDH^{+}99] Yves Bertot, Gilles Dowek, André Hirschowitz, Christine Paulin, and Laurent Théry, editors. Theorem proving in higher order logics: 12th International Conference, TPHOLs ’99, Nice, France, September 14–17, 1999: Proceedings, volume 1690 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66463-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1690; QA76.9.A96 T655 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1690.htm;
 http://www.springerlink.
com/content/978-3-540-
66463-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1690.

Beliczynski:2007:ANCa

[BDIR07a] Bartlomiej Beliczynski, Andrzej Dzielinski, Marcin Iwanowski, and Bernardete Ribeiro, editors. Adaptive and Natural Computing Algorithms: 8th International Conference, ICANNGA 2007, Warsaw, Poland, April 11–14, 2007, Proceedings, Part I, volume 4431 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71589-4 (print), 3-540-71618-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71618-1.

Beliczynski:2007:ANCb

[BDIR07b] Bartlomiej Beliczynski, Andrzej Dzielinski, Marcin Iwanowski, and Bernardete Ribeiro, editors. Adaptive and Natural Computing Algorithms: 8th International Conference, ICANNGA 2007, Warsaw, Poland, April 11–14, 2007, Proceedings, Part II, volume 4432 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71590-8 (print), 3-540-71629-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71629-7.

Balandin:2010:SSN

[BDK10a] Sergey Balandin, Roman Dunaytsev, and Yevgeni Koucheryavy, editors. Smart Spaces and Next Generation Wired/ Wireless Networking: Third Conference on Smart Spaces, ruSMART 2010, and 10th International Conference, NEW2AN 2010, St. Petersburg, Russia, August 23–25, 2010. Proceedings, volume 6294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14890-5 (print), 3-642-14891-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14891-0.

Bolduc:2010:MPC

[BDK10b] Claude Bolduc, Jules Desharnais, and Béchir Ktari, editors. Mathematics of Program Construction: 10th International Conference, MPC 2010, Québec City, Canada, June 21–23, 2010. Proceedings, volume 6120 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13320-7 (print), 3-642-13321-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13321-3.

Beckmann:2008:LTA

[BDL08] Arnold Beckmann, Costas Dimitracopoulos, and Benedikt Löwe, editors. Logic and Theory of Algorithms: 4th Conference on Computability in Europe, CiE 2008, Athens, Greece, June 15–20, 2008 Proceedings, volume 5028 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69405-6 (print), 3-540-69407-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69407-6.

Busemeyer:2012:QII

[BDLMM12] Jerome R. Busemeyer, François Dubois, Ariane Lambert-Mogiliansky, and Massimo Melucci, editors. Quantum Interaction: 6th International Symposium, QI 2012, Paris, France, June 27–29, 2012, Revised Selected Papers, volume 7620 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35658-3 (print), 3-642-35659-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35659-9.

Branke:2008:MOI

[BDMS08] Jürgen Branke, Kalyanmoy Deb, Kaisa Miettinen, and Roman Słowiński, editors. Multiobjective Optimization: Interactive and Evolutionary Approaches, volume 5252 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88907-8 (print), 3-540-88908-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88908-3.

Boujemaa:2008:AMR

[BDN08] Nozha Boujemaa, Marcin Detyniecki, and Andreas Nürnberger, editors. Adaptive Multimedial Retrieval: Retrieval, User, and Semantics: 5th International Workshop, AMR 2007, Paris, France, July 5–6, 2007 Revised Selected Papers, volume 4918 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79859-5 (print), 3-540-79860-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79860-6.

Barthe:2002:ASI

[BDPS02] Gilles Barthe, Peter Dybjer, Luís Pinto, and João Saraiva, editors. Applied Semantics: International Summer School, APPSEM 2000, Caminha, Portugal, September 9–15, 2000: Advanced Lectures, volume 2395 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44044-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .A68 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2395.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44044-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2395.

Beklemishev:2011:LLI

[BdQ11] Lev D. Beklemishev and Ruy de Queiroz, editors. Logic, Language, Information and Computation: 18th International Workshop, WoL-LIC 2011, Philadelphia, PA, USA. Proceedings, volume 6642 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20919-X (print), 3-642-20920-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20920-8.

Bergadano:1994:MLE

[BdR94]  Francesco  Bergadano

and Luc de Raedt, editors. Machine learning: ECML-94: European Conference on Machine Learning, Catania, Italy, April 6–8, 1994: proceedings, volume 784 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57868-4 (New York), 3-540-57868-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 1994. DM80.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0784.htm;
 http://www.springerlink.
com/content/978-0-387-

                                        57868-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=784.

Billington:1999:APN

[BDR99] Jonathan Billington, M. Diaz, and Grzegorz Rozenberg. Application of Petri nets to communication networks: advances in Petri nets, volume 1605 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65870-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 303 pp. LCCN QA76.9.D5 A67 1999.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1605.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65870-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1605.

Blass:2010:FLC

[BDR10] Andreas Blass, Nachum Dershowitz, and Wolfgang Reisig, editors. Fields of Logic and Computation: Essays Dedicated to Yuri Gurevich on the Occasion of His 70th Birthday, volume 6300 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15024-1 (print), 3-642-15025-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15025-8.

Boiten:2004:IFM

[BDS04] Eerke A. Boiten, John Derrick, and Graeme Smith, editors. Integrated Formal Methods: 4th International Conference, IFM 2004, Canterbury, UK, April 4–7, 2004, Proceedings, volume 2999 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21377-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2999.htm;
http://www.springerlink.
com/content/978-3-540-
21377-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2999;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96106.

Bertels:2009:ECS

[BDSW09] Koen Bertels, Nikitas Dimopoulos, Cristina Silvano, and Stephan Wong, editors. Embedded Computer Systems: Architectures, Modeling, and Simulation: 9th International Workshop, SAMOS 2009, Samos, Greece, July 20– 23, 2009. Proceedings, volume 5657 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03137-4 (print), 3-642-03138-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03138-0.

Braccini:1995:IAP

[BDV95] Carlo Braccini, Leila De-Floriani, and Gianni Vernazza, editors. Image analysis and processing: 8th International Conference, ICIAP ’95, San Remo, Italy, September 13–15, 1995: proceedings, volume 974 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60298-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .I552 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0974.htm;
 http://www.springerlink.
com/content/978-3-540-
60298-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=974.

Bubak:1997:RAP

[BDW97] Marian Bubak, J. J.

Dongarra, and Jerzy Waśniewski, editors. Recent advances in parallel virtual machine and message passing interface: 4th European PVM/MPI user’s group meeting Cracow, Poland, November 3–5, 1997: proceedings, volume 1332 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63697-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E973 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1332.htm;
 http://www.springerlink.
com/content/978-3-540-
63697-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1332.

Berekovic:2008:ECS

[BDW08] Mladen Bereković, Nikitas Dimopoulos, and Stephan Wong, editors. Embedded Computer Systems: Architectures, Modeling, and Simulation: 8th International Workshop, SAMOS 2008, Samos, Greece, July 21–24, 2008. Proceedings, volume 5114 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70549-X (print), 3-540-70550-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70550-5.

Bao:2004:PKC

[BDZ04] Feng Bao, Robert Deng, and Jianying Zhou, editors. Public Key Cryptography—PKC 2004: 7th International Workshop on Practice and Theory in Public Key Cryptography, Singapore, March 1–4, 2004: Proceedings, volume 2947 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21018-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I567 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2947.
htm;   http://www.springerlink.
com/content/978-3-540-
21018-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2947;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95631.

Bernardo:2008:FMC

[BDZ08] Marco Bernardo, Pierpaolo Degano, and Gianluigi Zavattaro, editors. Formal Methods for Computational Systems Biology: 8th International School on Formal Methods for the Design of Computer, Communication, and Software Systems, SFM 2008 Bertinoro, Italy, June 2– 7, 2008 Advanced Lectures, volume 5016 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68892-7 (print), 3-540-68894-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        68894-5.

Bauer:1974:CCA

[BE74] Friedrich Ludwig Bauer and J. Eickel. Compiler construction: an advanced course, volume 21 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 621 pp. LCCN QA76.6 .C6281 1976. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0021.htm;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=21.

Bauer:1976:CCA

[BE76] Friedrich Ludwig Bauer and J. Eickel. Compiler construction: an advanced course, volume 21 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, second edition, 1976. CODEN LNCSD9. ISBN 0-387-08046-5 (USA), 3-540-08046-5 (Berlin), 0-387-07542-9. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 638 pp. LCCN QA76.6 .C628 1976. URL http:

                                        //link.springer-ny.

com/link/service/series/
0558/tocs/t0021.htm;
http://www.springerlink.
com/content/978-0-387-
08046-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=21.

Becker:1987:WPP

[BE87] J. Becker and I. Eisele, editors. WOPPLOT 86: parallel processing: logic, organization, and technology: proceedings of a workshop, Neubiberg, Federal Republic of Germany, July 2– 4, 1986, volume 253 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18022-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .W6851 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0253.htm;
http://www.springerlink.
com/content/978-0-387-
18022-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=253.

Banzhaf:1995:EBC

[BE95] Wolfgang Banzhaf and Frank H. Eeckman. Evolution and biocomputation: computational models of evolution, volume 899 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59046-3, 0-387-59046-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 277 pp. LCCN QH455 .E92 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0899.htm;
 http://www.springerlink.
com/content/978-3-540-
59046-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=899.

Burke:2001:PTA

[BE01] Edmund Burke and Wilhelm Erben, editors. Practice and theory of automated timetabling III: Third International Conference, PATAT 2000, Konstanz, Germany, August 16– 18, 2000: Selected Papers, volume 2079 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42421-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2079; Q341 .P7 2000. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2079.
htm;   http://www.springerlink.
com/content/978-3-540-
42421-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2079.

Brandes:2005:NAM

[BE05] Ulrik Brandes and Thomas Erlebach, editors. Network analysis: methodological foundations, volume 3418 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24979-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA402 .N48 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24979-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3418;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106453.

Baldoni:2006:DALb

[BE06] Matteo Baldoni and Ulle Endriss, editors. Declarative Agent Languages and Technologies IV: 4th International Workshop, DALT 2006, Hakodate, Japan, May 8, 2006, Selected, Revised and Invited Papers, volume 4327 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68959-1 (print), 3-540-68961-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68961-4.

Bosnacki:2007:MCS

[BE07] Dragan Bošnački and Stefan Edelkamp, editors. Model Checking Software: 14th International SPIN Workshop, Berlin, Germany, July 1–3, 2007. Proceedings, volume 4595 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73369-8 (print), 3-540-73370-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73370-6.

Bello:2008:BSI

[BE08] Fernando Bello and P. J. Eddie Edwards, editors. Biomedical Simulation: 4th International Symposium, IS-BMS 2008, London, UK, July 7–8, 2008 Proceedings, volume 5104 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70520-1 (print), 3-540-70521-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70521-5.

Bach:2011:KAA

[BE11] Joscha Bach and Stefan Edelkamp, editors. KI 2011: Advances in Artificial Intelligence: 34th Annual German Conference on AI, Berlin, Germany, October 4–7,2011. Proceedings, volume 7006 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24454-8 (print), 3-642-24455-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        24455-1.

Becvar:1975:MFC

[Bec75] J. Becvar, editor. Mathematical foundations of computer science 1975, 4th Symposium, Marianske Lazne, September 1–5, 1975, volume 32 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CO-DEN LNCSD9. ISBN 0-387-07389-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.S4 .S89 1975.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0032.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        07389-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=32.

“Organized by the Mathematical Institute of the Czechoslovak Academy of Sciences and is cosponsored by the International Federation for Information Processing (IFIP).”.

Becvar:1979:MFC

[Bec79] J. Becvar, editor. Mathematical foundations of computer science, 1979: proceedings, 8th Symposium, Olomouc, Czechoslovakia, September 3–7, 1979, volume 74 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CODEN LNCSD9. ISBN 0-387-09526-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S9194 1979. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0074.htm;
 http://www.springerlink.
com/content/978-0-387-
09526-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=74.

Beckert:2005:ARA

[Bec05] Bernhard Beckert, editor. Automated reasoning with analytic tableaux and related methods: international conference, TABLEAUX 2005, Koblenz, Germany, September 14– 17, 2005: proceedings, volume 3702 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28931-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T33 2005.

URL http://www.springerlink.
com/content/978-3-540-

                                        28931-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3702.

Beech:1986:CUI

[Bee86] David Beech. Concepts in user interfaces: a reference model for command and response languages, volume 234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16791-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 115 pp. LCCN QA267.A1 L43 no.234. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0234.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        16791-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=234.

Beer:1989:CDP

[Bee89] Joachim Beer. Concepts, design, and performance analysis of a parallel Prolog machine, volume 404 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-52053-8 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 128 pp. LCCN QA76.73.P76 B44 1989. DM28.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0404.htm;
 http://www.springerlink.
com/content/978-0-387-
52053-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=404.

Beetz:2000:CRP

[Bee00] Michael Beetz. Concurrent Reactive Plans: Anticipating and Forestalling Execution Failures, volume 1772 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67241-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 213 pp. LCCN TJ211.495. B44 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1772.
htm;   http://www.springerlink.
com/content/978-3-540-
67241-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1772.

Beetz:2002:PBC

[Bee02] Michael Beetz. Plan-Based Control of Robotic Agents: Improving the Capabilities of Autonomous Robots, volume 2554 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00335-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 191 pp. LCCN TJ211.35 .B45 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2554.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        00335-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2554.

Also available via the

World Wide Web.

Broersma:2008:GTC

[BEFP08] Hajo Broersma, Thomas Erlebach, Tom Friedetzky, and Daniel Paulusma, editors. Graph-Theoretic Concepts in Computer Science: 34th International Workshop, WG 2008, Durham, UK, June 30 – July 2, 2008. Revised Papers, volume 5344 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92247-4 (print), 3-540-92248-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92248-3.

Boulicaut:2004:MLE

[BEG04] Jean-François Boulicaut, Floriana Esposito, and Fosca Giannotti, editors. Machine Learning: ECML 2004: 15th European Conference on Machine Learning, Pisa, Italy, September 20– 24, 2004: Proceedings, volume 3201 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23105-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3201.htm;
 http://www.springerlink.
com/content/978-3-540-
23105-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3201;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b100702.

Boulicaut:2004:KDD

[BEGP04] Jean-François Boulicaut, Floriana Esposito, Fosca Giannotti, and Dino Pedreschi, editors. Knowledge Discovery in Databases: PKDD 2004 8th European Conference on Principles and Practice of Knowledge Discovery in Databases, Pisa, Italy, September 20–24, 2004: Proceedings, volume 3202 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23108-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9 D3 P16 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3202.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23108-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3202;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100704.

Bauer:1985:MPCb

[BEH^{+}85] Friedrich Ludwig Bauer, H. Ehler, A. Horsch, B. Möller, H. Partsch, O. Paukner, and P. Pepper, editors. The Munich Project CIP: The Program Transformation System CIP-S, volume 292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CODEN LNCSD9. ISBN 0-387-18779-0 (New York), 3-540-18779-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 522 pp. LCCN QA76.6 .M86 1985 v. 1– 2 (c1985–1987). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0292.
htm;   http://www.springerlink.
com/content/978-0-387-
18779-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=292.

Behnke:2003:HNN

[Beh03] Sven Behnke, editor. Hierarchical Neural Networks for Image Interpretation, volume 2766 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40722-7. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 224 pp. LCCN TA1634 .B44 2003. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t2766.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40722-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2766;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11963.

Benz:2011:LGE

[BEJvR11] Anton Benz, Christian Ebert, Gerhard Jäger, and Robert van Rooij, editors. Language, Games, and Evolution: Trends in Current Research on Language and Game Theory, volume 6207 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18005-1 (print), 3-642-18006-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18006-4.

Brauer:2002:FNC

[BEKS02] Wilfried Brauer, H. Ehrig, J. Karhumäki, and A. K. Salomaa, editors. Formal and Natural Computing: essays dedicated to Grzegorz Rozenberg, volume 2300 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43190-X. ISSN 0302-9743 (print), 1611-3349 (electronic). xxxvi + 429 pp. LCCN QA76 .F625 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2300.htm;
 http://www.springerlink.
com/content/978-3-540-
43190-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2300.

Bellare:2000:ACC

[Bel00] Mihir Bellare, editor. Advances in Cryptology— CRYPTO 2000: 20th Annual International Cryptology Conference, Santa Barbara, California, USA, August 20– 24, 2000. Proceedings, volume 1880 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67907-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1880.

                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67907-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1880.

Bellahsene:2003:DXT

[Bel03] Zohra Bellahsène, editor. Database and XML Technologies: First International XML Database Symposium, XSym 2003, Berlin, Germany, September 8, 2003: Proceedings, volume 2824 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20055-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2824.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20055-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2824;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13480.

Becker:1991:PLE

[BEM91] J. Becker, I. Eisele, and F. W. Mundemann, editors. Parallelism, learning, evolution: Workshop on Evolutionary Models and Strategies, Neubiberg, Germany, March 10–11, 1989: Workshop on Parallel Processing: Logic, Organization, and Technology, WOPPLOT 89, Wildbad Kreuth, Germany, July 24–28, 1989: proceedings, volume 565 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-55027-5 (Berlin), 0-387-55027-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .W685 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0565.htm;
 http://www.springerlink.
com/content/978-3-540-
55027-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=565.

Bourque:2006:CGR

[BEM06] Guillaume Bourque and Nadia El-Mabrouk, editors. Comparative Genomics: RECOMB 2006 International Workshop, RCG 2006 Montreal, Canada, September 24– 26, 2006 Proceedings, volume 4205 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44529-3 (print), 3-540-44530-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        44530-2.

Benhamou:2006:PPC

[Ben06] Frédéric Benhamou, editor. Principles and Practice of Constraint Programming — CP 2006: 12th International Conference, CP 2006, Nantes, France, September 25–29, 2006. Proceedings, volume 4204 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46267-8 (print), 3-540-46268-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46268-2.

Baldoni:2006:DALa

[BEOT06] Matteo Baldoni, Ulle Endriss, Andrea Omicini, and Paolo Torroni, editors. Declarative Agent Languages and Technologies III: Third International Workshop, DALT 2005, Utrecht, The Netherlands, July 25, 2005, Selected and Revised Papers, volume 3904 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33106-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33106-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3904.

Belz:2004:NLG

[BEP04] Anja Belz, Roger Evans, and Paul Piwek, editors. Natural Language Generation: Third International Conference, INLG 2004, Brockenhurst, UK, July 14–16, 2004: Proceedings, volume 3123 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22340-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 N389 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3123.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        22340-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3123;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98634.

Bertino:1996:CSE

[Ber96] Elisa Bertino, editor. Computer security, ES-ORICS 96: 4th European Symposium on Research in Computer Security, Rome, Italy, September 25–27, 1996: proceedings, volume 1146 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61770-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E967 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1146.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61770-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1146.

Bert:1998:BRA

[Ber98] Didier Bert, editor. B’98: recent advances in the development and use of the B method: Second International B Conference, Montpellier, France, April 22– 24, 1998: Proceedings, volume 1393 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64405-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1393. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1393.htm;
 http://www.springerlink.
com/content/978-3-540-
64405-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1393.

Bergmann:1999:DIC

[Ber99] R. Bergmann. Developing industrial case-based reasoning applications: the INRECA methodology, volume 1612 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66182-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xx + 188 pp. LCCN QA76.76.E95 D485 1999 Accompanying CD-Rom shelved in Reserves. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1612.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66182-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1612.

Bertino:2000:EOO

[Ber00] Elisa Bertino, editor. ECOOP 2000 — objectoriented programming: 14th European Conference, Sophia Antipolis and Cannes, France, June 12–16, 2000: Proceedings, volume 1850 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67660-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E24 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1850.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67660-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1850.

Bergmann:2002:EMF

[Ber02a] R. (Ralph) Bergmann. Experience Management: Foundations, Development Methodology, and Internet-Based Applications, volume 2432 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44191-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xx + 393 pp. LCCN HD30.2 .B465 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2432.htm;
 http://www.springerlink.
com/content/978-3-540-
44191-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2432.

Bert:2002:ZFS

[Ber02b] Didier Bert, editor. ZB 2002: formal specification and development in Z and B: 2nd International Conference of B and Z Users, Grenoble, France, January 23–25, 2002: Proceedings, volume 2272 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43166-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.B155 I58 2002.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t2272.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        43166-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2272.

Bertossi:2003:SDS

[Ber03] Leopoldo Bertossi, editor. Semantics in databases: Second International Workshop, Dagstuhl Castle, Germany, January 7–12, 2001: Revised Papers, volume 2582 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00957-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S438 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2582.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00957-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2582.

Bergler:2008:AAI

[Ber08] Sabine Bergler, editor. Advances in Artificial Intelligence: 21st Conference of the Canadian Society for Computational Studies of Intelligence, Canadian AI 2008 Windsor, Canada, May 28–30, 2008 Proceedings, volume 5032 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68821-8 (print), 3-540-68825-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68825-9.

Berger:2010:RCM

[Ber10] Bonnie Berger, editor.

Research in Computational Molecular Biology: 14th Annual International Conference, RECOMB 2010, Lisbon, Portugal, April 25– 28, 2010. Proceedings, volume 6044 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12682-0 (print), 3-642-12683-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12683-3.

Berthold:2012:BKD

[Ber12] Michael R. Berthold, editor. Bisociative Knowledge Discovery: An Introduction to Concept, Algorithms, Tools, and Applications, volume 7250 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31829-0 (print), 3-642-31830-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31830-6.

Best:1993:CIC

[Bes93] Eike Best, editor. CON-CUR ’93: 4th International Conference on Concurrency Theory, Hildesheim, Germany, August 993: proceedings, volume 715 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57208-2 (New York), 3-540-57208-2 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0715.htm;
 http://www.springerlink.
com/content/978-0-387-

57208-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=715.

Bessiere:2007:PPC

[Bes07] Christian Bessière, editor. Principles and Practice of Constraint Programming – CP 2007: 13th International Conference, CP 2007, Providence, RI, USA, September 23–27, 2007. Proceedings, volume 4741 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74969-1 (print), 3-540-74970-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74970-7.

Beth:1983:CPW

[Bet83] Thomas Beth, editor.

Cryptography: proceedings of the Workshop on Cryptography, Burg Feuerstein, Germany, March 29–April 2, 1982, volume 149 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CODEN LNCSD9. ISBN 0-387-11993-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z102.5 .C78 1983. DM43.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0149.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        11993-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=149.

Boral:1989:DMS

[BF89] H. Boral and P. Faudemay, editors. Database machines: sixth International Workshop, IWDM ’89, Deauville, France, June 19–21, 1989: proceedings, volume 368 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51324-8 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D35761 1989. DM52.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0368.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51324-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=368.

“Papers presented at the Sixth International Workshop on Database Machines”–Pref.

Bergstra:1991:AMI

[BF91] J. A. Bergstra and L. M. G. Feijs, editors. Algebraic methods II: theory, tools, and applications, volume 490 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-53912-3 (Berlin), 0-387-53912-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 A45 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0490.htm;
http://www.springerlink.
com/content/978-3-540-
53912-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=490.

“Originates from a workshop organized by ES-PRIT project no 432 METEOR in Mierlo, the Netherlands, in September 1989”–Pref.

Biundo:2000:RAA

[BF00] Susanne Biundo and Maria Fox, editors. Recent advances in AI planning: 5th European Conference on Planning, ECP’99, Durham, UK, September 8–10, 1999:

Proceedings, volume 1809 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67866-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .E37 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1809.htm;
 http://www.springerlink.
com/content/978-3-540-
67866-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1809.

Bussler:2004:AIM

[BF04] Christoph Bussler and Dieter Fensel, editors. Artificial Intelligence: Methodology, Systems, and Applications, volume 3192 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22959-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I554 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3192.htm;
 http://www.springerlink.
com/content/978-3-540-
22959-9;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3192;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100218.

Borwein:2006:MKM

[BF06] Jonathan M. Borwein and William M. Farmer, editors. Mathematical Knowledge Management: 5th International Conference, MKM 2006, Wokingham, UK, August 11–12, 2006. Proceedings, volume 4108 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37104-4 (print), 3-540-37106-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37106-9.

Barthe:2008:TGC

[BF08] Gilles Barthe and Cédric Fournet, editors. Trustworthy Global Computing: Third Symposium, TGC 2007, Sophia-Antipolis, France, November 5–6, 2007, Revised Selected Papers, volume 4912 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78662-7 (print), 3-540-78663-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78663-4.

Bergel:2009:SCI

[BF09] Alexandre Bergel and Johan Fabry, editors. Software Composition: 8th International Conference, SC 2009, Zurich, Switzerland, July 2– 3, 2009. Proceedings, volume 5634 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02654-0 (print), 3-642-02655-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02655-3.

Bottaci:2010:TPR

[BF10] Leonardo Bottaci and Gordon Fraser, editors. Testing — Practice and Research Techniques: 5th International Academic and Industrial Conference, TAIC PART 2010, Windsor, UK, September 3–5, 2010. Proceedings, volume 6303 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15584-7 (print), 3-642-15585-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15585-7.

Berry:2011:REF

[BF11] Daniel Berry and Xavier Franch, editors. Requirements Engineering: Foundation for Software Quality: 17th International Working Conference, REFSQ 2011, Essen, Germany, March 28–30, 2011. Proceedings, volume 6606 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19857-0 (print), 3-642-19858-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19858-8.

Beringer:2012:ITP

[BF12] Lennart Beringer and Amy Felty, editors. Interactive Theorem Proving: Third International Conference, ITP 2012, Princeton, NJ, USA, August 13–15, 2012. Proceedings, volume 7406 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32346-4 (print), 3-642-32347-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32347-8.

Berekovic:2011:ACS

[BFBS11] Mladen Berekovic, William Fornaciari, Uwe Brinkschulte, and Cristina Silvano, editors. Architecture of Computing Systems — ARCS 2011: 24th International Conference, Como, Italy, February 24–25, 2011. Proceedings, volume 6566 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19136-3 (print), 3-642-19137-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19137-4.

Butz:2009:SGI

[BFC^{+}09] Andreas Butz, Brian Fisher, Marc Christie, Antonio Krüger, Patrick Olivier, and Roberto Therón, editors. Smart Graphics: 10th International Symposium, SG 2009, Salamanca, Spain, May 28–30, 2009. Proceedings, volume 5531 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02114-X (print), 3-642-02115-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02115-2.

Barringer:2010:RVF

[BFF^{+}10] Howard Barringer, Ylies Falcone, Bernd Finkbeiner, Klaus Havelund, Insup Lee, Gordon Pace, Grigore Roşu, Oleg Sokolsky, and Nikolai Tillmann, editors. Runtime Verification: First International Conference, RV 2010, St. Julians, Malta, November 1–4, 2010. Proceedings, volume 6418 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16611-3 (print), 3-642-16612-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16612-9.

Bielikova:2012:STP

[BFG^{+}12] Mária Bieliková, Gerhard Friedrich, Georg Gottlob, Stefan Katzenbeisser, and György Turán, editors. SOF-SEM 2012: Theory and Practice of Computer Science: 38th Conference on Current Trends in Theory and Practice of Computer Science, Špindleruv Mlýn, Czech Republic, January 21–27, 2012. Proceedings, volume 7147 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27659-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27659-0;   http://www.
                                        springerlink.com/content/
                                        978-3-642-27659-0/.

Bernhaupt:2010:HCS

[BFGL10] Regina Bernhaupt, Peter Forbrig, Jan Gulliksen, and Marta Lárusdóttir, editors. Human-Centred Software Engineering: Third International Conference, HCSE 2010, Reykjavik, Iceland, October 14–15, 2010. Proceedings, volume 6409 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16487-0 (print), 3-642-16488-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16488-0.

Banatre:2005:UPP

[BFGM05] Jean-Pierre Banâtre, Pascal Fradet, Jean-Louis Giavitto, and Olivier Michel, editors. Unconventional Programming Paradigms: International Workshop UPP 2004, Le Mont Saint Michel, France, September 15–17, 2004, Revised Selected and Invited Papers, volume 3566 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27884-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .U66 2004. URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        27884-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3566.

Bowen:1998:ZZF

[BFH98] J. P. Bowen, Andreas Fett, and Michael G. Hinchey, editors. ZUM’98: the Z formal specification notation: 11th International Conference of Z Users, Berlin, Germany, September 24–26, 1998: Proceedings, volume 1493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65070-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1493.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1493.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65070-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1493.

Baresi:2007:WEI

[BFH07] Luciano Baresi, Piero Fraternali, and Geert-Jan Houben, editors. Web Engineering: 7th International Conference, ICWE 2007 Como, Italy, July 16–20, 2007 Proceedings, volume 4607 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73596-8 (print), 3-540-73597-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73597-7.

Bonchi:2009:PST

[BFJM09] Francesco Bonchi, Elena Ferrari, Wei Jiang, and Bradley Malin, editors. Privacy, Security, and Trust in KDD: Second ACM SIGKDD International Workshop, PinKDD 2008, Las Vegas, NV, USA, August 24, 2008, Revised Selected Papers, volume 5456 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01717-7 (print), 3-642-01718-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01718-6.

Brauer:1997:FCS

[BFJV97] Wilfried Brauer, C. Freksa, Matthias Jantzen, and R. (Rudiger) Valk, editors. Foundations of computer science: potential — theory — cognition, volume 1337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63746-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.F6374 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1337.
htm;   http://www.springerlink.
com/content/978-3-540-
63746-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1337.

Butz:2007:SGI

[BFK^{+}07] Andreas Butz, Brian Fisher, Antonio Krüger, Patrick Olivier, and Shigeru Owada, editors. Smart Graphics: 8th International Symposium, SG 2007, Kyoto, Japan, June 25–27, 2007. Proceedings, volume 4569 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73213-6 (print), 3-540-73214-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73214-3.

Butz:2008:SGI

[BFK^{+}08] Andreas Butz, Brian Fisher, Antonio Krüger, Patrick Olivier, and Marc Christie, editors. Smart Graphics: 9th International Symposium, SG 2008, Rennes, France, August 27–29, 2008. Proceedings, volume 5166 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85410-X (print), 3-540-85412-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85412-8.

Butz:2005:SGI

[BFKO05] Andreas Butz, Brian Fisher, Antonio Krüger, and Patrick Olivier, editors. Smart graphics: 5th international symposium, SG 2005, Frauenwörth Cloister, Germany, August 22– 24, 2005: proceedings, volume 3638 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28179-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .I592 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28179-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3638.

Butz:2006:SGI

[BFKO06] Andreas Butz, Brian Fisher, Antonio Krüger, and Patrick Olivier, editors. Smart Graphics: 6th International Symposium, SG 2006, Vancouver, Canada, July 23–25, 2006. Proceedings, volume 4073 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36293-2 (print), 3-540-36295-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        36295-1.

Bouge:1996:EPPa

[BFMR96a] L. (Luc) Bougé, Pierre Fraigniaud, Anne Mignotte, and Yves Robert, editors. Euro-Par’96: Parallel Processing: Second International Euro-Par Conference, Lyon, France, August 26–29, 1996: Proceedings. Volume 1, volume 1123 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61626-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I554 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1123.
htm;   http://www.springerlink.
com/content/978-3-540-
61626-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1123.

Bouge:1996:EPPb

[BFMR96b] L. (Luc) Bougé, Pierre Fraigniaud, Anne Mignotte, and Yves Robert, editors. Euro-Par’96: Parallel Processing: Second International Euro-Par Conference, Lyon, France, August 26–29, 1996: Proceedings. Volume 2, volume 1124 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61627-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I554 1996. URL

http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t1124.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61627-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1124.

Brodal:2001:AEI

[BFMS01] Gerth Stolting Brodal, Daniele Frigioni, and Alberto Marchetti-Spaccamela, editors. Algorithm engineering: 5th International Workshop, WAE 2001, Aarhus, Denmark, August 28– 31, 2001: Proceedings, volume 2141 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42500-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I59 2001; QA267.A1 L43 no.2141. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2141.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42500-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2141.

Bonchi:2008:PST

[BFMS08] Francesco Bonchi, Elena Ferrari, Bradley Malin, and Yücel Saygin, editors. Privacy, Security, and Trust in KDD: First ACM SIGKDD International Workshop, PinKDD 2007, San Jose, CA, USA, August 12, 2007, Revised Selected Papers, volume 4890 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78477-2 (print), 3-540-78478-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78478-4.

Bussler:2004:WSB

[BFOY04] Christoph Bussler, Dieter Fensel, Maria E. Orlowska, and Jian Yang, editors. Web Services, E-Business, and the Semantic Web: Second International Workshop, WES 2003, Klagenfurt, Austria, June 16–17, 2003: Revised Selected Papers, volume 3095 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22396-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN x + 145. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3095.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22396-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3095;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98797.

Biundo:2004:KAA

[BFP04] Susanne Biundo, Thom Frühwirth, and Günther Palm, editors. KI 2004: advances in artificial intelligence: 27th Annual German Conference on AI, KI 2004, Ulm, Germany, September 20–24, 2004: proceedings, volume 3238 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23166-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .K485 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        23166-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3238;

http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100351.

Beth:1992:PKC

[BFS92] Thomas Beth, M. Frisch, and Gustavus J. Simmons, editors. Publickey cryptography: state of the art and future directions: E.I.S.S workshop, Oberwolfach, Germany, July 3–6, 1991: final report, volume 578 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55215-4 (Berlin), 0-387-55215-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 P83 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0578.htm;
http://www.springerlink.
com/content/978-3-540-
55215-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=578.

Bohme:2010:IHI

[BFSN10] Rainer Böhme, Philip W. L. Fong, and Reihaneh Safavi-Naini, editors. Information Hiding: 12th International Conference, IH 2010, Calgary, AB, Canada, June 28–30, 2010, Revised Selected Papers, volume 6387 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16434-X (print), 3-642-16435-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16435-4.

Bouge:2008:EPW

[BFT^{+}08] Luc Bougé, Martti Forsell, Jesper Larsson Träff, Achim Streit, Wolfgang Ziegler, Michael Alexander, and Stephen Childs, editors. Euro-Par 2007 Workshops: Parallel Processing: HPPC 2007, UNICORE Summit 2007, and VHPC 2007, Rennes, France, August 28–31, 2007, Revised Selected Papers, volume 4854 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78472-1 (print), 3-540-78474-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
78474-6.

Brennan:2010:MAC

[BFvdM10] Rob Brennan, Joel Fleck II, and Sven van der Meer, editors. Modelling Autonomic Communication Environments: 5th IEEE International Workshop, MACE 2010, Niagara Falls, Canada, October 28, 2010. Proceedings, volume 6473 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16835-3 (print), 3-642-16836-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16836-9.

Baeten:1991:CIC

[BG91a] J. C. M. Baeten and J. F. Groote, editors. CONCUR ’91: 2nd International Conference on Concurrency Theory, Amsterdam, The Netherlands, August 26– 29, 1991, Proceedings, volume 527 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-54430-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.527.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0527.
htm;   http://www.springerlink.
com/content/978-0-387-
54430-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=527.

“CONCUR ’91 is the second international conference on concurrency theory, organised in association with the NFI project Transfer”–Pref.

Beaumont:1991:PEL

[BG91b] A. Beaumont and G. Gupta, editors. Parallel execution of logic programs: ICLP ’91 pre-conference workshop, Paris, June 24, 1991: proceedings, volume 569 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-55038-0 (New York), 3-540-55038-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I54 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0569.htm;
 http://www.springerlink.
com/content/978-0-387-
55038-1;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=569.

“These papers were presented at the workshop on parallel logic programming held in Paris on June 24th as part of the 8th International Conference on Logic Programming”–Pref.

Barrett:1993:IPM

[BG93a] Harrison H. Barrett and Arthur F. Gmitro, editors. Information processing in medical imaging: 13th International Conference, IPMI ’93 Flagstaff, Arizona, USA, June 14–18, 1993 proceedings, volume 687 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56800-X (Berlin), 0-387-56800-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN RC78.7.D53 I54 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0687.htm;
 http://www.springerlink.
com/content/978-3-540-
56800-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=687.

Bezem:1993:TLC

[BG93b] M. Bezem and J. F. Groote, editors. Typed lambda calculi and applications: International Conference on Typed Lam[b]da Calculi and Applications, TLCA ’93, March 16–18, 1993, Utrecht, the Netherlands: proceedings, volume 664 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56517-5 (Berlin), 0-387-56517-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.5 .I58 1993.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0664.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        56517-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=664.

Backhouse:2003:GPA

[BG03a] Roland Backhouse and Jeremy Gibbons, editors. Generic Programming: Advanced Lectures, volume 2793 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20194-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6245 .G46 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2793.
htm;   http://www.springerlink.
com/content/978-3-540-
20194-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2793;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/

b12027. Based on material presented at a summer school on Generic Programming held August 2002 at the University of Oxford.

Bigun:2003:IAS

[BG03b] Josef Bigun and Tomas Gustavsson, editors. Image Analysis: 13th Scandinavian Conference, SCIA 2003, Halmstad, Sweden, June 29– July 2, 2003: Proceedings, volume 2749 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40601-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .S27 2003. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t2749.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40601-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2749.

Boyd:2005:ISP

[BG05] Colin Boyd and Juan M. González Nieto, editors. Information security and privacy: 10th Australasian conference, ACISP 2005, Brisbane, Australia, July 4–6, 2005: proceedings, volume 3574 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26547-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 A278 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26547-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3574.

Biere:2006:TAS

[BG06] Armin Biere and Carla P. Gomes, editors. Theory and Applications of Satisfiability Testing — SAT 2006: 9th International Conference, Seattle, WA, USA, August 12–15, 2006. Proceedings, volume 4121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37206-7 (print), 3-540-37207-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37207-3.

Borisov:2007:PET

[BG07a] Nikita Borisov and Philippe Golle, editors. Privacy Enhancing Technologies: 7th International Symposium, PET 2007 Ottawa, Canada, June 20– 22, 2007 Revised Selected Papers, volume 4776 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75550-0 (print), 3-540-75551-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75551-7.

Bshouty:2007:LTA

[BG07b] Nader H. Bshouty and Claudio Gentile, editors. Learning Theory: 20th Annual Conference on Learning Theory, COLT 2007, San Diego, CA, USA; June 13–15, 2007. Proceedings, volume 4539 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72925-9 (print), 3-540-72927-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72927-3.

Borisov:2008:PET

[BG08] Nikita Borisov and Ian Goldberg, editors. Privacy Enhancing Technologies: 8th International Symposium, PETS 2008 Leuven, Belgium, July 23–25, 2008 Proceedings, volume 5134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70629-1 (print), 3-540-70630-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70630-4.

Bartolini:2009:IMS

[BG09] Claudio Bartolini and Luciano Paschoal Gaspary, editors. Integrated Management of Systems, Services, Processes and People in IT: 20th IFIP/IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2009, Venice, Italy, October 27–28, 2009. Proceedings, volume 5841 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04988-5 (print), 3-642-04989-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04989-7.

Babar:2010:SAE

[BG10a] Muhammad Ali Babar and Ian Gorton, editors. Software Architecture: 4th European Conference, ECSA 2010, Copenhagen, Denmark, August 23–26, 2010. Proceedings, volume 6285 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15113-2 (print), 3-642-15114-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15114-9.

Boldi:2010:FAI

[BG10b] Paolo Boldi and Luisa Gargano, editors. Fun with Algorithms: 5th International Conference, FUN 2010, Ischia, Italy, June 2–4, 2010. Proceedings, volume 6099 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13121-2 (print), 3-642-13122-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13122-6.

Benferhat:2011:SUM

[BG11a] Salem Benferhat and John Grant, editors. Scalable Uncertainty Management: 5th International Conference, SUM 2011, Dayton, OH, USA, October 10–13, 2011. Proceedings, volume 6929 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23962-5 (print), 3-642-23963-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23963-2.

Bettstetter:2011:SOS

[BG11b] Christian Bettstetter and Carlos Gershenson, editors. Self-Organizing Systems: 5th International Workshop, IW-SOS 2011, Karlsruhe, Germany, February 23– 24. 2011. Proceedings, volume 6557 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19166-5 (print), 3-642-19167-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19167-1.

Bezhanishvili:2012:LLC

[BG12a] Nick Bezhanishvili and Valentin Goranko, editors. Lectures on Logic and Computation: ESS-LLI 2010 Copenhagen, Denmark, August 2010, ESSLLI 2011, Ljubljana, Slovenia, August 2011, Selected Lecture Notes, volume 7388 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31484-8 (print), 3-642-31485-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31485-8.

Bikakis:2012:RWR

[BG12b] Antonis Bikakis and Adrian Giurca, editors. Rules on the Web: Research and Applications: 6th International Symposium, RuleML 2012, Montpellier, France, August 27–29, 2012. Proceedings, volume 7438 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32688-9 (print), 3-642-32689-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32689-9.

Berthold:2005:CLS

[BGD^{+}05] Michael R. Berthold, Robert Glen, Kay Diederichs, Oliver Kohlbacher, and Ingrid Fischer, editors. Computational life sciences: first international symposium, CompLife 2005, Konstanz, Germany, September 25– 27, 2005: proceedings, volume 3695 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29104-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .C66 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29104-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3695.

Berthold:2006:CLS

[BGF06] Michael R. Berthold, Robert C. Glen, and Ingrid Fischer, editors. Computational Life Sciences II: Second International Symposium, CompLife 2006, Cambridge, UK, September 27–29, 2006. Proceedings, volume 4216 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45767-4 (print), 3-540-45768-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45768-8.

Bao:2012:WAI

[BGG^{+}12] Zhifeng Bao, Yunjun Gao, Yu Gu, Longjiang Guo, Yingshu Li, Jiaheng Lu, Zujie Ren, Chaokun Wang, and Xiao Zhang, editors. Web-Age Information Management: WAIM 2012 International Workshops: GDMM, IWSN, MDSP, USDM, and XMLDM Harbin, China, August 18–20, 2012 Proceedings, volume 7419 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33049-5 (print), 3-642-33050-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33050-6.

Barthe:2006:CAS

[BGH06a] Gilles Barthe, Benjamin Gregoire, and Marieke Huisman, editors. Construction and analysis of Safe, Secure, and Interoperable Smart Devices: Second International Workshop, CAS-SIS 2005, Nice, France, March 8–11, 2005, Revised Selected Papers, volume 3956 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33689-3 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33689-1.

Butterfield:2006:IAF

[BGH06b] Andrew Butterfield, Clemens Grelck, and Frank Huch, editors. Implementation and Application of Functional Languages: 17th International Workshop, IFL 2005, Dublin, Ireland, September 19–21, 2005, Revised Selected Papers, volume 4015 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69174-X (print), 3-540-69175-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69175-4.

Backhouse:2007:DGP

[BGHJ07] Roland Backhouse, Jeremy Gibbons, Ralf Hinze, and Johan Jeuring, editors. Datatype-Generic Programming: International Spring School, SSDGP 2006, Nottingham, UK, April 24– 27, 2006, Revised Lectures, volume 4719 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76785-1 (print), 3-540-76786-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76786-2.

Bresciani:2005:AOI

[BGHS05] Paolo Bresciani, Paolo Giorgini, and Brian Henderson-Sellers, editors. Agent-Oriented Information Systems II: 6th International Bi-Conference Workshop, AOIS 2004 Riga, Latvia, June 8, 2004 and New York, NY, USA, July 20, 2004. Revised Selected Papers, volume 3508 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25911-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25911-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3508;
 http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b136434.

Bosse:2011:MAB

[BGJ11] Tibor Bosse, Armando Geller, and Catholijn M. Jonker, editors. Multi-Agent-Based Simulation XI: International Workshop, MABS 2010, Toronto, Canada, May 11, 2010, Revised Selected Papers, volume 6532 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18344-1 (print), 3-642-18345-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18345-4.

Barros:2012:BPM

[BGK12] Alistair Barros, Avigdor Gal, and Ekkart Kindler, editors. Business Process Management: 10th International Conference, BPM 2012, Tallinn, Estonia, September 3–6, 2012. Proceedings, volume 7481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32884-9 (print), 3-642-32885-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32885-5.

Bouzeghoub:2004:SNW

[BGKS04] Mokrane Bouzeghoub, Carole Goble, Vipul Kashyap, and Stefano Spaccapietra, editors. Semantics of a Networked World: Semantics for Grid Databases, First International IFIP Conference, ICSNW 2004, Paris, France, June 17–19, 2004, Revised Selected Papers, volume 3226 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23609-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3226.htm;
 http://www.springerlink.
com/content/978-3-540-
23609-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3226;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b102069.

Bellovin:2008:ACN

[BGKY08] Steven M. Bellovin, Rosario Gennaro, Angelos Keromytis, and Moti Yung, editors. Applied Cryptography and Network Security: 6th International Conference, ACNS 2008, New York, NY, USA, June 3–6, 2008. Proceedings, volume 5037 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68913-3 (print), 3-540-68914-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68914-0.

Bosch:2009:LLC

[BGL09] Peter Bosch, David Gabelaia, and Jérôme Lang, editors. Logic, Language, and Computation: 7th International Tbilisi Symposium on Logic, Language, and Computation, TbiLLC 2007, Tbilisi, Georgia, October 1–5, 2007. Revised Selected Papers, volume 5422 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00664-7 (print), 3-642-00665-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00665-4.

Baral:2005:LPN

[BGLT05] Chitta Baral, Gianluigi Greco, Nicola Leone, and Giorgio Terracina, editors. Logic Programming and Nonmonotonic Reasoning: 8th International Conference, LPNMR 2005, Diamante, Italy, September 5–8, 2005: proceedings, volume 3662 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28538-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28538-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3662.

Borger:1994:CSL

[BGM94] E. Borger, Yuri Gurevich, and K. Meinke, editors. Computer science logic: 7th workshop, CSL ’93, Swansea, United Kingdom, September 13– 17, 1993: Selected Papers, volume 832 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58277-0 (Berlin), 0-387-58277-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .W64 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0832.htm;
 http://www.springerlink.
com/content/978-3-540-
58277-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=832.

Buntine:2009:MLKa

[BGMST09a] Wray Buntine, Marko Grobelnik, Dunja Mladenić, and John Shawe-Taylor, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2009, Bled, Slovenia, September 7–11, 2009, Proceedings, Part I, volume 5781 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04179-5 (print), 3-642-04180-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04180-8.

Buntine:2009:MLKb

[BGMST09b] Wray Buntine, Marko Grobelnik, Dunja Mladenić, and John Shawe-Taylor, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2009, Bled, Slovenia, September 7–11, 2009, Proceedings, Part II, volume 5782 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04173-6 (print), 3-642-04174-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04174-7.

Bongiovanni:2000:ACI

[BGP00] Giancarlo Bongiovanni, G. Gambosi, and Rosella Petreschi, editors. Algorithms and Complexity: 4th Italian Conference, CIAC 2000, Rome, Italy, March 1–3, 2000: Proceedings, volume 1767 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67159-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .I885 2000; QA267.A1 L43 no.1767. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1767.htm;
 http://www.springerlink.
com/content/978-3-540-
67159-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1767.

Bassiliades:2008:RRI

[BGP08] Nick Bassiliades, Guido Governatori, and Adrian Paschke, editors. Rule Representation, Interchange and Reasoning on the Web: International Symposium, RuleML 2008, Orlando, FL, USA, October 30– 31, 2008. Proceedings, volume 5321 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88807-1 (print), 3-540-88808-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88808-6.

Bassiliades:2011:RBR

[BGP11] Nick Bassiliades, Guido Governatori, and Adrian Paschke, editors. Rule-Based Reasoning, Programming, and Applications: 5th International Symposium, RuleML 2011 — Europe, Barcelona, Spain, July 19–21, 2011. Proceedings, volume 6826 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22545-4 (print), 3-642-22546-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22546-8.

Borodovsky:2010:BRA

[BGPR10] Mark Borodovsky, Johann Peter Gogarten, Teresa M. Przytycka, and Sanguthevar Rajasekaran, editors. Bioinformatics Research and Applications: 6th International Symposium, IS-BRA 2010, Storrs, CT, USA, May 23–26, 2010. Proceedings, volume 6053 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13077-1 (print), 3-642-13078-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13078-6.

Berre:2010:FIF

[BGPTF10] Arne J. Berre, Asunción Gómez-Pérez, Kurt Tutschku, and Dieter Fensel, editors. Future Internet

— FIS 2010: Third Future Internet Symposium, Berlin, Germany, September 20–22, 2010. Proceedings, volume 6369 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15876-5 (print), 3-642-15877-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15877-3.

Bohlen:2005:GTE

[BGPW05] Michael Böhlen, Johann Gamper, Wolfgang Polasek, and Maria A. Wimmer, editors. Egovernment: Towards Electronic Democracy: International Conference, TCGOV 2005 Bolzano, Italy, March 2–4, 2005 Proceedings, volume 3416 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25016-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN JF1525.A8 T34 2005. URL http:

//www.springerlink.
com/content/978-3-540-
25016-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3416;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106542.

Borger:2003:ASM

[BGR03] E. Börger, Angelo Gargantini, and Elvinia Riccobene, editors. Abstract state machines 2003: advances in theory and practice: 10th International Workshop, ASM 2003, Taormina, Italy, March 3–7, 2003: Proceedings, volume 2589 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00624-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A25 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2589.htm;
http://www.springerlink.
com/content/978-3-540-
00624-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2589.

Brinkschulte:2008:STE

[BGR08] Uwe Brinkschulte, Tony Givargis, and Stefano Russo, editors. Software Technologies for Embedded and Ubiquitous Systems: 6th IFIP WG 10.2 International Workshop, SEUS 2008, Anacarpi, Capri Island, Italy, October 1–3, 2008 Proceedings, volume 5287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87784-3 (print), 3-540-87785-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87785-1.

Blanken:2003:ISX

[BGS^{+}03] Henk M. Blanken, Torsten Grabs, Hans-Jörg Schek, Ralf Schenkel, and Gerhard Weikum, editors. Intelligent Search on XML Data: Applications, Languages, Models, Implementations, and Benchmarks, volume 2818 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40768-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xvii + 319 pp. LCCN QA76.9.D3. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2818.htm;
 http://www.springerlink.
com/content/978-3-540-
40768-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2818;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13249.

Biryukov:2011:SAC

[BGS11] Alex Biryukov, Guang Gong, and Douglas R. Stinson, editors. Selected Areas in Cryptography: 17th International Workshop, SAC 2010, Waterloo, Ontario, Canada, August 12–13, 2010, Revised Selected Papers, volume 6544 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19573-3 (print), 3-642-19574-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19574-7.

Blumenthal:1994:HCI

[BGU94] Brad Blumenthal, Iu. M.

Gornostaev, and C. Unger, editors. Human-computer interaction: 4th international conference, EWHCI ’94, St Petersburg, Russia, August 2– 5, 1994: Selected Papers, volume 876 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58648-2 (New York), 3-540-58648-2 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I58 1994. DM52.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0876.
htm;   http://www.springerlink.
com/content/978-0-387-
58648-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=876.

Blumenthal:1995:HCI

[BGU95] Brad Blumenthal, Iu. M.

Gornostaev, and C. Unger, editors. Human-computer interaction: 5th international conference, EWHCI ’95, Moscow, Russia, July 3–7, 1995: Selected Papers, volume 1015 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60614-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 H86 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1015.htm;
 http://www.springerlink.
com/content/978-3-540-
60614-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1015.

Bauer:2001:UMI

[BGV01] Mathias Bauer, Piotr J. Gmytrasiewicz, and Julita Vassileva, editors. User modeling 2001: 8th International Conference, UM 2001, Sonthofen, Germany, July 13–17, 2001: Proceedings, volume 2109 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42325-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I599 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2109.htm;
http://www.springerlink.
com/content/978-3-540-
42325-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2109.

Barak:1993:MDO

[BGW93] Amnon Barak, Shai Guday, and Richard G. Wheeler. The MOSIX distributed operating system: load balancing for UNIX, volume 672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56663-5 (Berlin), 0-387-56663-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 221 pp. LCCN QA76.76.O63 B363 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0672.
htm;   http://www.springerlink.
com/content/978-3-540-
56663-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=672.

Buttyan:2006:SPA

[BGW06] Levente Buttyán, Virgil D. Gligor, and Dirk Westhoff, editors. Security and Privacy in Ad-Hoc and Sensor Networks: Third European Workshop, ESAS 2006, Hamburg, Germany, September 20–21, 2006, Revised Selected Papers, volume 4357 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69172-3 (print), 3-540-69173-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69173-0.

Briot:1996:OBP

[BGY96] Jean-Pierre Briot, Jean-Marc Geib, and Akinori Yonezawa, editors. Object-based parallel and distributed computation: France-Japan Workshop, OBPDC ’95, Tokyo, Japan, June 21– 23, 1995: Selected Papers, volume 1107 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61487-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .F7 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1107.htm;
http://www.springerlink.
com/content/978-3-540-
61487-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1107.

Blaser:1977:ISP

[BH77] A. Blaser and Clemens Hackl, editors. Interactive systems: proceedings, 6th Informatik Symposium, IBM Germany, Bad Homburg v. d. H., September 1976, volume 49 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CODEN LNCSD9. ISBN 0-387-08141-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.I58I531 1976.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0049.
htm;   http://www.springerlink.
com/content/978-0-387-
08141-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=49.

English or German.

Biskup:1992:DTI

[BH92] J. Biskup and Richard Hull, editors. Database theory — ICDT ’92: 4th International Conference, Berlin, Germany, October 14–16, 1992: proceedings, volume 646 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-56039-4, 0-387-56039-4 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I557 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0646.
htm;   http://www.springerlink.
com/content/978-3-540-
56039-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=646.

Spine title: ICDT ’92, database theory.

Bowen:1995:ZFS

[BH95] J. P. Bowen and Michael G. Hinchey, editors. ZUM ’95, the Z formal specification notation: 9th International Conference of Z Users, Limerick, Ireland, September 7–9, 1995: proceedings, volume 967 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60271-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.Z2 I58 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0967.
htm;   http://www.springerlink.
com/content/978-3-540-
60271-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=967.

Brause:2000:MDA

[BH00] Rudiger W. Brause and Ernst Hanisch, editors. Medical Data Analysis: First International Symposium, ISMDA 2000, Frankfurt, Germany, September 29–30, 2000. Proceedings, volume 1933 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41089-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1933. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1933.htm;
http://www.springerlink.
com/content/978-3-540-
41089-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1933.

Borriello:2002:UUC

[BH02] Gaetano Borriello and Lars Erik Holmquist, editors. UbiComp 2002:

Ubiquitous Computing: 4th International Conference, Göteborg, Sweden, September 29–October 1, 2002. Proceedings, volume 2498 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44267-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5915 .U25 2002.

URL http://link.springer.
de/link/service/series/
0558/tocs/t2498.htm;
 http://www.springerlink.
com/content/978-3-540-
44267-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2498.

Also available via the World Wide Web.

Bozanis:2005:AIP

[BH05] Panayiotis Bozanis and Elias N. Houstis, editors. Advances in Informatics: 10th Panhellenic Conference on Informatics, PCI 2005, Volas, Greece, November 11–13, 2005. Proceedings, volume 3746 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29673-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29673-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3746.

Baier:2006:CCT

[BH06a] Christel Baier and Holger Hermanns, editors. CONCUR 2006 — Concurrency Theory: 17th International Conference, CONCUR 2006, Bonn, Germany, August 27–30, 2006. Proceedings, volume 4137 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37376-4 (print), 3-540-37377-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37377-3.

Baresi:2006:FAS

[BH06b] Luciano Baresi and Reiko Heckel, editors. Fundamental Approaches to Software Engineering: 9th International Conference, FASE 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 27–28, 2006. Proceedings, volume 3922 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33093-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33093-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3922.

Bell:2006:FEI

[BH06c] David A. Bell and Jun Hong, editors. Flexible and Efficient Information Handling: 23rd British National Conference on Databases, BN-COD 23, Belfast, Northern Ireland, UK, July 18–20, 2006. Proceedings, volume 4042 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35969-9 (print), 3-540-35971-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35971-5.

Bussler:2006:BPM

[BH06d] Christoph Bussler and Armin Haller, editors. Business Process Management Workshops: BPM 2005 International Workshops, BPI, BPD, ENEI, BPRM, WSCOBPM, BPS, Nancy, France, September 5, 2005. Revised Selected Papers, volume 3812 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32595-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32595-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3812.

Bernardo:2007:FMP

[BH07] Marco Bernardo and Jane Hillston, editors. Formal Methods for Performance Evaluation: 7th International School on Formal Methods for the Design of Computer, Communication, and Software Systems, SFM 2007, Bertinoro, Italy, May 28–June 2, 2007, Advanced Lectures, volume 4486 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72482-6 (print), 3-540-72522-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72522-0.

Beckert:2008:TPS

[BH08] Bernhard Beckert and Reiner Hähnle, editors. Tests and Proofs: Second International Conference, TAP 2008, Prato, Italy, April 9–11, 2008. Proceedings, volume 4966 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79123-X (print), 3-540-79124-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79124-9.

Barthe:2010:VMC

[BH10] Gilles Barthe and Manuel Hermenegildo, editors. Verification, Model Checking, and Abstract Interpretation: 11th International Conference, VMCAI 2010, Madrid, Spain, January 17–19, 2010. Proceedings, volume 5944 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11318-4 (print), 3-642-11319-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11319-2.

Bultan:2011:ATV

[BH11] Tevfik Bultan and Pao-Ann Hsiung, editors. Automated Technology for Verification and Analysis: 9th International Symposium, ATVA 2011, Taipei, Taiwan, October 11–14, 2011. Proceedings, volume 6996 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24371-1 (print), 3-642-24372-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24372-1.

Bhalla:1995:ISD

[Bha95] Subhash C. Bhalla, editor. Information systems and data management:

6th international conference, CISMOD ’95, Bombay, India, November 15–17, 1995: proceedings, volume 1006 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60584-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55863 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1006.htm;
 http://www.springerlink.
com/content/978-3-540-
60584-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1006.

Bhalla:2000:DNI

[Bha00] Subhash Bhalla, editor. Databases in Networked Information Systems: International Workshop DNIS 2000, Aizu, Japan, December 4–6, 2000. Proceedings, volume 1966 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41395-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D586 2000. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t1966.htm;
http://www.springerlink.
com/content/978-3-540-
41395-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1966.

Bhalla:2002:DNI

[Bha02] Subhash Bhalla, editor. Databases in networked information systems: Second International Workshop, DNIS 2002, Aizu, Japan, December 16–18, 2002: Proceedings, volume 2544 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00264-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D586 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2544.htm;
http://www.springerlink.
com/content/978-3-540-
00264-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2544.

Also available via the World Wide Web.

Bhalla:2005:DNI

[Bha05] Subhash Bhalla, editor. Databases in Networked Information Systems: 4th International Workshop, DNIS 2005, Aizu-Wakamatsu, Japan, March 28–30, 2005: Proceedings, volume 3433 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25361-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D586 2005.

URL http://www.springerlink.
com/content/978-3-540-
25361-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3433;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106805.

Bhalla:2007:DNI

[Bha07] Subhash Bhalla, editor. Databases in Networked Information Systems: 5th International Workshop, DNIS 2007, Aizu-Wakamatsu, Japan, October 17–19, 2007. Proceedings, volume 4777 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75511-X (print), 3-540-75512-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75512-8.

Bezivin:1987:EEC

[BHCL87] J. Bézivin, J.-M. Hullot, P. Cointe, and H. Lieberman, editors. ECOOP ’87, European Conference on Object-Oriented Programming: Paris, France, June 15–17, 1987: proceedings, volume 276 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18353-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E973 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0276.htm;
http://www.springerlink.
com/content/978-0-387-
18353-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=276.

Beetz:2002:APB

[BHGP02] Michael Beetz, Joachim Hertzberg, Malik Ghallab, and Martha E. Pollack, editors. Advances in plan-based control of robotic agents: international seminar, Dagstuhl Castle, Germany, October 21–26, 2001: Revised Papers, volume 2466 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00168-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ211.35 .A28 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2466.htm;
 http://www.springerlink.
com/content/978-3-540-
00168-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2466.

Brunnstein:1974:RFR

[BHH74] K. Brunnstein, Klaus Haefner, and Wolfgang Handler, editors. RGU ’74: Fachtagung, Rechner-Gestützter Unterricht, Hamburg, 12.- 14 August, 1974. (German) [RGU ’74 Technical Conference, Computeraided instruction, Hamburg, August 12–14, 1974], volume 17 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 0-387-06907-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.5 .R36. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0017.htm;
 http://www.springerlink.
com/content/978-0-387-
06907-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=17.

Papers in English or German.

Baier:2004:VSS

[BHH^{+}04a] Christel Baier, Boudewijn R.

Haverkort, Holger Hermanns, Joost-Pieter Katoen, and Markus Siegle, editors. Validation of Stochastic Systems: A Guide to Current Research, volume 2925 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22265-0. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 465 pp. LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2925.htm;
 http://www.springerlink.
com/content/978-3-540-
22265-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2925;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98484.

Barillot:2004:MICa

[BHH04b] Christian Barillot, David R. Haynor, and Pierre Hellier, editors. Medical Image Computing and Computer-Assisted Intervention—MICCAI 2004: 7th International Conference, Saint-Malo, France, September 26– 29, 2004, Proceedings, Part I, volume 3216 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22976-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3216.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22976-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3216;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100265.

Barillot:2004:MICb

[BHH04c] Christian Barillot, David R. Haynor, and Pierre Hellier, editors. Medical Image Computing and Computer-Assisted Intervention—MICCAI 2004: 7th International Conference, Saint-Malo, France, September 26– 29, 2004: Proceedings, Part II, volume 3217 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22977-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3217.htm;
 http://www.springerlink.
com/content/978-3-540-
22977-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3217;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100270.

Bremer:2006:KDL

[BHH^{+}06] Eric G. Bremer, Jörg Hakenberg, Eui-Hong (Sam) Han, Daniel Berrar, and Werner Dubitzky, editors. Knowledge Discovery in Life Science Literature: PAKDD 2006 International Workshop, KDLL 2006, Singapore, April 9, 2006. Proceedings, volume 3886 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32809-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32809-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3886.

Bui:2008:IAM

[BHH08] The Duy Bui, Tuong Vinh Ho, and Quang Thuy Ha, editors. Intelligent Agents and Multi-Agent Systems: 11th Pacific Rim International Conference on Multi-Agents, PRIMA 2008, Hanoi, Vietnam, December 15–16, 2008. Proceedings, volume 5357 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89673-2 (print), 3-540-89674-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89674-6.

Bobaru:2011:NFM

[BHHJ11] Mihaela Bobaru, Klaus Havelund, Gerard J. Holzmann, and Rajeev Joshi, editors. NASA Formal Methods: Third International Symposium, NFM 2011, Pasadena, CA, USA, April 18– 20, 2011. Proceedings, volume 6617 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20397-3 (print), 3-642-20398-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20398-5.

Bechhofer:2008:SWR

[BHHK08] Sean Bechhofer, Manfred Hauswirth, Jörg Hoffmann, and Manolis Koubarakis, editors. The Semantic Web: Research and Applications: 5th European Semantic Web Conference, ESWC 2008, Tenerife, Canary Islands, Spain, June 1–5, 2008 Proceedings, volume 5021 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68233-3 (print), 3-540-68234-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68234-9.

Bose:2011:CLS

[BHJ^{+}11] Jürgen W. Böse, Hao Hu, Carlos Jahn, Xiaoning Shi, Robert Stahlbock, and Stefan Voß, editors. Computational Logistics: Second International Conference, ICCL 2011, Hamburg, Germany, September 19– 22, 2011. Proceedings, volume 6971 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24263-4 (print), 3-642-24264-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24264-9.

Barni:2006:IHI

[BHJKPG06] Mauro Barni, Jordi Herrera-Joancomartí, Stefan Katzenbeisser, and Fernando Pérez-González, editors. Information Hiding: 7th International Workshop, IH 2005, Barcelona, Spain, June 6–8, 2005. Revised Selected Papers, volume 3727 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-29039-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29039-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3727.

Brueckner:2007:ESO

[BHJY07] Sven A. Brueckner, Salima Hassas, Márk Jelasity, and Daniel Yamins, editors. Engineering Self-Organising Systems: 4th International Workshop, ESOA 2006, Hakodate, Japan, May 9, 2006, Revised and Invited Papers, volume 4335 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69867-1 (print), 3-540-69868-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69868-5.

Brinksma:2001:LFM

[BHK01] Ed Brinksma, Holger Hermanns, and Joost-Pieter Katoen, editors. Lectures on formal methods and performance analysis: first EEF/ Euro Summer School on Trends in Computer Science, Berg en Dal, The Netherlands, July 3–7, 2000: revised lectures, volume 2090 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42479-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 E34 2000; QA267.A1 L43 no.2090. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2090.htm;
 http://www.springerlink.
com/content/978-3-540-
42479-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2090.

Barner:2011:HSV

[BHKR11] Sharon Barner, Ian Harris, Daniel Kroening, and Orna Raz, editors. Hardware and Software: Verification and Testing: 6th International Haifa Verification Conference, HVC 2010, Haifa, Israel, October 4–7, 2010. Revised Selected Papers, volume 6504 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19582-2 (print), 3-642-19583-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19583-9.

Bjorner:1990:VVZ

[BHL90] D. Bjorner, C. A. R.

Hoare, and Hans Langmaack, editors. VDM ’90: VDM and Z: formal methods in software development: Third International Symposium of VDM Europe, Kiel, FRG, April 17–21, 1990: proceedings, volume 428 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52513-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 V36 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0428.htm;
http://www.springerlink.
com/content/978-0-387-
52513-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=428.

Bakker:2003:IVR

[BHL^{+}03] Erwin M. Bakker, Thomas S.

Huang, Michael S. Lew, Nicu Sebe, and Xiang (Sean) Zhou, editors. Image and Video Retrieval: Second International Conference, CIVR 2003 Urbana-Champaign, IL, USA, July 24–25, 2003 Proceedings, volume 2728. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40634-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .C49 2003. URL

http://www.springerlink.
com/content/978-3-540-
40634-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2728.

Bouguettaya:2011:WIS

[BHL11] Athman Bouguettaya, Manfred Hauswirth, and Ling Liu, editors. Web Information System Engineering — WISE 2011: 12th International Conference, Sydney, Australia, October 13–14, 2011. Proceedings, volume 6997 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24433-5 (print), 3-642-24434-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24434-6.

Butler:2007:FMS

[BHLP07] Michael Butler, Michael G.

Hinchey, and María M. Larrondo-Petrie, editors. Formal Methods and Software Engineering: 9th International Conference on Formal Engineering Methods, ICFEM 2007, Boca Raton, FL, USA, November 14–15, 2007. Proceedings, volume 4789 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76648-0 (print), 3-540-76650-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76650-6.

Brim:2007:FMA

[BHLvdP07] Luboš Brim, Boudewijn Haverkort, Martin Leucker, and Jaco van de Pol, editors. Formal Methods: Applications and Technology: 11th International Workshop, FMICS 2006 and 5th International Workshop PDMC 2006, Bonn, Germany, August 26–27, and August 31, 2006, Revised Selected Papers, volume 4346 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70951-7 (print), 3-540-70952-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70952-7.

Bussler:2002:WSB

[BHM^{+}02] Christoph Bussler, Rick Hull, Sheila McIlraith, Maria E. Orlowska, Barbara Pernici, and Jian Yang, editors. Web Services, E-Business, and the Semantic Web: CAiSE 2002 International Workshop, WES 2002, Toronto, Canada, May 27–28, 2002. Revised Papers, volume 2512 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00198-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815 .W47 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2512.htm;
 http://www.springerlink.
com/content/978-3-540-

00198-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2512.

Also available via the World Wide Web.

Bry:2003:PPS

[BHM03] François Bry, Nicola Henze, and Jan Małuszyński, editors. Principles and Practice of Semantic Web Reasoning: International Workshop, PP-SWR 2003, Mumbai, India, December 8, 2003: Proceedings, volume 2901 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20582-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815 .P67 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2901.
htm;   http://www.springerlink.
com/content/978-3-540-
20582-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2901;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94439.

Berendt:2004:WMW

[BHM^{+}04] Bettina Berendt, Andreas Hotho, Dunja Mladenic, Maarten van Someren, Myra Spiliopoulou, and Gerd Stumme, editors. Web Mining: From Web to Semantic Web: First European Web Mining Forum, EWMF 2003, Cavtat-Dubrovnik, Croatia, September 22, 2003: Invited and Selected Revised Papers, volume 3209 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23258-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3209.htm;
 http://www.springerlink.
com/content/978-3-540-
23258-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3209;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100615.

Berendt:2007:WSW

[BHMS07] Bettina Berendt, Andreas Hotho, Dunja Mladenic, and Giovanni Semeraro, editors. From Web to Social Web: Discovering and Deploying User and Content Profiles: Workshop on Web Mining, WebMine 2006, Berlin, Germany, September 18, 2006. Revised Selected and Invited Papers, volume 4737 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74950-0 (print), 3-540-74951-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74951-6.

Brewka:1997:KAA

[BHN97] Gerhard Brewka, Christopher Habel, and Bernhard Nebel, editors. KI-97, advances in artificial intelligence: 21st Annual German Conference on Artificial Intelligence, Freiburg, Germany, September 9– 12, 1997: proceedings, volume 1303 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63493-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G466 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1303.

                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63493-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1303.

Baumgartner:1995:TPA

[BHP95] Peter Baumgartner, Reiner Hahnle, and Joachim Posegga, editors. Theorem proving with analytic tableaux and related methods: 4th International Workshop, TABLEAUX ’95, Schloss Rheinfels, St. Goar, Germany, May 7–10, 1995: proceedings, volume 918 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59338-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T43 1995.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0918.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        59338-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=918.

Bringas:2010:DESa

[BHQ10a] Pablo García Bringas, Abdelkader Hameurlain, and Gerald Quirchmayr, editors. Database and Expert Systems Applications: 21st International Conference, DEXA 2010, Bilbao, Spain, August 30 — September 3, 2010, Proceedings, Part I, volume 6261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15363-1 (print), 3-642-15364-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15364-8.

Bringas:2010:DESb

[BHQ10b] Pablo García Bringas, Abdelkader Hameurlain, and Gerald Quirchmayr, editors. Database and Expert Systems Applications: 21th International Conference, DEXA 2010, Bilbao, Spain, August 30 — September 3, 2010, Proceedings, Part II, volume 6262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15250-3 (print), 3-642-15251-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15251-1.

Borger:1984:LMD

[BHR84] E. Borger, G. Hasenjaeger, and D. Rodding, editors. Logic and machines: decision problems and complexity: Proceedings of the Symposium “Rekursive Kombinatorik” held from May 23–28, 1983 at the Institut für Mathematische Logik und Frundlagenforschung der Universität Munster/Westfalen, volume 171 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-13331-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .S98 1983. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0171.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        13331-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=171.

Blasius:1990:STA

[BHR90] K. H. (Karl Hans) Bläsius, U. (Ulrich)

Hedtstück, and Claus-Rainer Rollinger, editors. Sorts and types in artificial intelligence: workshop, Eringerfeld, FRG, April 24–26, 1989: proceedings, volume 418 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52337-5 (New York), 3-540-52337-5 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .S632 1990. DM47.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0418.htm;
 http://www.springerlink.
com/content/978-0-387-
52337-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=418.

“Workshop held in Geseke, West Germany … sponsored and organized by IBM Germany”–Introd.

Bellahsene:2009:DXT

[BHRU09] Zohra Bellahsène, Ela Hunt, Michael Rys, and Rainer Unland, editors. Database and XML Technologies: 6th International XML Database Symposium, XSym 2009, Lyon, France, August 24, 2009. Proceedings, volume 5679 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03554-X (print), 3-642-03555-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03555-5.

Bertossi:2005:IT

[BHS05] Leopoldo Bertossi, Anthony Hunter, and Torsten Schaub, editors. Inconsistency Tolerance, volume 3300 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24260-0. ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24260-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=volume&id=
                                        doi:10.1007/b104925.

Beckert:2007:VOO

[BHS07] Bernhard Beckert, Reiner Hähnle, and Peter H. Schmitt, editors. Verification of Object-Oriented Software. The KeY Approach, volume 4334 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-68977-X (print), 3-540-69061-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-

69061-0. Foreword by K. Rustan M. Leino.

Bowen:1997:ZFS

[BHT97] J. P. (Jonathan Peter) Bowen, Michael G. (Michael Gerard) Hinchey, and David Till, editors. ZUM ’97, the Z formal specification notation: 10th International Conference of Z Users, Reading, UK, April 3–4, 1997: proceedings, volume 1212 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62717-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.Z2 I58 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1212.
htm;   http://www.springerlink.
com/content/978-3-540-
62717-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1212.

Bohme:2003:IIC

[BHU03] Thomas Böhme, Gerhard Heyer, and Herwig Unger, editors. Innovative Internet Community Systems: Third International Workshop: Revised Papers/IICS 2003, Leipzig, Germany, June 19–21, 2003, volume 2877 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20436-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I52 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2877.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20436-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2877;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b93936.

Bravo:2011:AAL

[BHV11] José Bravo, Ramón Hervás, and Vladimir Villarreal, editors. Ambient Assisted Living: Third International Workshop, IWAAL 2011, Held at IWANN 2011, Torremolinos-Málaga, Spain, June 8–10, 2011. Proceedings, volume 6693 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21302-2 (print), 3-642-21303-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21303-8.

Biehl:2009:SBC

[BHVV09] Michael Biehl, Barbara Hammer, Michel Verleysen, and Thomas Villmann, editors. Similarity-Based Clustering: Recent Developments and Biomedical Applications, volume 5400 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01804-1 (print), 3-642-01805-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01805-3.

Bohling:1973:EFA

[BI73] Karl Heinz Bohling and Klaus Indermark, editors. 1. [i.e., Erste] Fachtagung über Automatentheorie und Formale Sprachen, Bonn, 9.–

12. Juli 1973. (German) [First Technical Conference on the Theory of Automata and Formal Languages], volume 2 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1973. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA 267 F119 1973. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0002.
htm;   http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&

volume=2. At head of title: GI-Gesellschaft für Informatik e.V. English, French, or German.

Bernardo:2003:FMS

[BI03] Marco Bernardo and Paola Inverardi, editors. Formal Methods for Software Architectures: Third International School on Formal Methods for the Design of Computer, Communication, and Software Systems—Software Architectures, SFM 2003, Bertinoro, Italy, September 2003: advanced lectures, volume 2804 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20083-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.754 .F67 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2804.htm;
 http://www.springerlink.
com/content/978-3-540-
20083-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2804;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13225.

Bomarius:2004:PFS

[BI04] Frank Bomarius and Hajimu Iida, editors. Product Focused Software Process Improvement: 5th International Conference, PROFES 2004, Kansai Science City, Japan, April 5–8, 2004: Proceedings, volume 3009 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21421-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .P7585 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3009.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21421-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3009;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96726.

Bernardo:2011:FME

[BI11] Marco Bernardo and Valérie Issarny, editors. Formal Methods for Eternal Networked Software Systems: 11th International School on Formal Methods for the Design of Computer, Communication and Software Systems, SFM 2011, Bertinoro, Italy, June 13– 18, 2011. Advanced Lectures, volume 6659 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21454-1 (print), 3-642-21455-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21455-4.

Biham:1997:FSE

[Bih97] Eli Biham, editor. Fast software encryption: 4th International Workshop, FSE ’97, Haifa, Israel, January 20–22, 1997: proceedings, volume 1267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63247-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25F77 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1267.
htm;   http://www.springerlink.
com/content/978-3-540-
63247-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1267.

Biham:2003:ACE

[Bih03] Eli Biham, editor. Advances in cryptology — EUROCRYPT 2003: International Conference on the Theory and Applications of Cryptographic Techniques, Warsaw, Poland, May 4–8, 2003: Proceedings, volume 2656 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-14039-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 2003. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t2656.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        14039-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2656.

Also available via the

World Wide Web.

Bertrand:2001:DIG

[BIK01] Gilles Bertrand, Atsushi Imiya, and Reinhard Klette, editors. Digital and Image Geometry: Advanced Lectures, volume 2243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43079-2. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 453 pp. LCCN TA1637 .D513 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2243.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43079-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2243.

Bilardi:1997:SIS

[Bil97] G. Bilardi, editor. Solving irregularly structured problems in parallel: 4th International Symposium, IRREGULAR ’97, Paderborn, Germany, June 12–13, 1997: proceedings, volume 1253 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63138-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5675 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1253.htm;
 http://www.springerlink.
com/content/978-3-540-
63138-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1253.

Bilardi:1998:AEA

[Bil98] G. (Gianfranco) Bilardi, editor. Algorithms — ESA ’98: 6th annual European symposium, Venice, Italy, August 24–26, 1998: Proceedings, volume 1461 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64848-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1461.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        64848-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1461.

Bouma:2012:NLP

[BIMW12] Gosse Bouma, Ashwin Ittoo, Elisabeth Métais, and Hans Wortmann, editors. Natural Language Processing and Information Systems: 17th International Conference on Applications of Natural Language to Information Systems, NLDB 2012, Groningen, The Netherlands, June 26–28, 2012. Proceedings, volume 7337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31177-6 (print), 3-642-31178-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31178-9.

Biryukov:2007:FSE

[Bir07] Alex Biryukov, editor. Fast software encryption: 14th international workshop, FSE 2007, Luxembourg, Luxembourg, March 26–28, 2007, revised selected papers, volume 4593 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-74617-X (softcover). LCCN QA76.9.A25 F77 2007. URL http:

//www.springerlink.
com/openurl.asp?genre=
book\%26isbn=978-3-
540-74617-1.

Birkedal:2012:FSS

[Bir12] Lars Birkedal, editor.

Foundations of Software Science and Computational Structures: 15th International Conference, FOSSACS 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7213 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28728-X (print), 3-642-28729-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28729-9.

Beigl:2005:UUC

[BIRT05] Michael Beigl, Stephen Intille, Jun Rekimoto, and Hideyuki Tokuda, editors. UbiComp 2005: ubiquitous computing: 7th international conference, UbiComp 2005, Tokyo, Japan, September 11–14, 2005: proceedings, volume 3660 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28760-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I584 2005. URL http:

//www.springerlink.
com/content/978-3-540-
28760-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3660.

Biskup:1988:MSM

[Bis88] J. Biskup, editor. MFDBS 87: 1st Symposium on Mathematical Fundamentals of Database Systems, Dresden, GDR, January 19–23, 1987: proceedings, volume 305 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 3-540-19121-6 (Berlin), 0-387-19121-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S961 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0305.htm;

http://www.springerlink.
com/content/978-3-540-
19121-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=305.

Bishop:2002:CDI

[Bis02] J. M. Bishop, editor. Component Deployment: IFIP/ACM Working Conference, CD 2002, Berlin, Germany, June 20–21, 2002. Proceedings, volume 2370 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43847-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C66 I355 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2370.
htm;   http://www.springerlink.
com/content/978-3-540-
43847-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2370.

Bistarelli:2004:SSC

[Bis04] Stefano Bistarelli. Semirings for Soft Constraint Solving and Programming, volume 2962 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21181-0. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 279 pp. LCCN QA76.612 .B57 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2962.htm;
 http://www.springerlink.
com/content/978-3-540-
21181-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2962;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95712.

Bittanti:1988:SRM

[Bit88] Sergio Bittanti. Software reliability modelling and identification, volume 341 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-50695-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 209 pp. LCCN QA76.76.R44 S661 1988.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0341.
htm;   http://www.springerlink.
com/content/978-0-387-
50695-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=341.

Bjorner:1978:VDM

[BJ78] D. Bjorner and C. B.

Jones. The Vienna Development Method: the Meta-language, volume 61 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1978. CODEN LNCSD9. ISBN 0-387-08766-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 382 pp. LCCN QA76.7 .V53. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0061.htm;
 http://www.springerlink.
com/content/978-0-387-
08766-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=61.

Bekic:1984:PLT

[BJ84] H. (Hans) Bekic and C. B. Jones. Programming languages and their definition, volume 177 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CODEN LNCSD9. ISBN 0-387-13378-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xxxii + 254 pp. LCCN QA76.7 .B45 1984. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0177.htm;

                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        13378-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=177.

Bibel:1986:MMS

[BJ86a] W. Bibel and K. P. Jantke, editors. Mathematical methods of specification and synthesis of software systems ’85: proceedings of the international spring school, Wendisch-Rietz, GDR, April 22–26, 1985, volume 215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CODEN LNCSD9. ISBN 0-387-16444-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M37 M371 1986. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0215.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        16444-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=215.

Bibel:1986:FAI

[BJ86b] W. Bibel and Ph. Jorrand. Fundamentals of artificial intelligence: an advanced course, volume 232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CODEN LNCSD9. ISBN 0-387-16782-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). v + 313 pp. LCCN Q335 .F861 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0232.htm;
 http://www.springerlink.
com/content/978-0-387-
16782-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=232.

Broy:1995:KML

[BJ95] M. Broy and Stefan Jähnichen. KORSO: methods, languages, and tools for the construction of correct software: final report, volume 1009 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60589-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 449 pp. LCCN QA76.76.D47 K67 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1009.
htm;   http://www.springerlink.
com/content/978-3-540-
60589-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1009.

Boldt:2001:AII

[BJ01a] Oliver Boldt and Helmut Jürgensen, editors. Automata implementation: 4th International Workshop on Implementing Automata, WIA’99, Potsdam, Germany, July 17–19, 1999: Revised Papers, volume 2214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42812-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2214; QA267 .I5787 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2214.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42812-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2214.

Boulton:2001:TPH

[BJ01b] Richard J. Boulton and Paul B. Jackson, editors. Theorem proving in higher order logics: 14th International Conference, TPHOLs 2001, Edinburgh, Scotland UK, September 3–6, 2001: Proceedings, volume 2152 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42525-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I577 2001; QA267.A1 L43 no.2152. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2152.htm;
 http://www.springerlink.
com/content/978-3-540-
42525-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2152.

Brazdil:2001:PAI

[BJ01c] Pavel B. Brazdil and Alipio Jorge, editors. Progress in Artificial Intelligence Knowledge Extraction, Multi-agent Systems, Logic Programming, and Constraint Solving: 10th Portuguese Conference on Artificial Intelligence, EPIA 2001, Porto, Portugal, December 17–20, 2001. Proceedings, volume 2258 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-43030-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P67 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2258.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43030-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2258.

Butler:2001:GCB

[BJ01d] G. Butler and Stan Jarzabek, editors. Generative and componentbased software engineering: Second International Symposium, GCSE 2000, Erfurt, Germany, October 9–12, 2000: Revised Papers, volume 2177 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42578-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .G418 2000; QA267.A1 L43 no.2177. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2177.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42578-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2177.

Ball:2006:CAV

[BJ06] Thomas Ball and Robert B. Jones, editors. Computer Aided Verification: 18th International Conference, CAV 2006, Seattle, WA, USA, August 17–20, 2006. Proceedings, volume 4144 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37406-X (print), 3-540-37411-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37411-4.

Bonsangue:2007:FMO

[BJ07] Marcello M. Bonsangue and Einar Broch Johnsen, editors. Formal Methods for Open Object-Based Distributed Systems: 9th IFIP WG 6.1 International Conference, FMOODS 2007, Paphos, Cyprus, June 6–8, 2007. Proceedings, volume 4468 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72919-4 (print), 3-540-72952-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        72952-5.

Bampis:2010:AOA

[BJ10] Evripidis Bampis and Klaus Jansen, editors. Approximation and Online Algorithms: 7th International Workshop,WAOA 2009, Copenhagen Denmark, September 10– 11, 2009. Revised Papers, volume 5893 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12449-6 (print), 3-642-12450-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12450-1.

Bonato:2012:AMW

[BJ12a] Anthony Bonato and Jeannette Janssen, editors. Algorithms and Models for the Web Graph: 9th International Workshop, WAW 2012, Halifax, NS, Canada, June 22–23, 2012. Proceedings, volume 7323 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30540-7 (print), 3-642-30541-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30541-2.

Brucker:2012:TPI

[BJ12b] Achim D. Brucker and Jacques Julliand, editors. Tests and Proofs: 6th International Conference, TAP 2012, Prague, Czech Republic, May 31 — June 1, 2012. Proceedings, volume 7305 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30472-9 (print), 3-642-30473-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30473-6.

Broy:2005:MBT

[BJK^{+}05] Manfred Broy, Bengt Jonsson, Joost-Pieter Katoen, Martin Leucker, and Alexander Pretschner, editors. Model-Based Testing of Reactive Systems: Advanced Lectures, volume 3472 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26278-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.T48 M644 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        26278-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3472.

Bampis:2006:EAO

[BJK06] Evripidis Bampis, Klaus Jansen, and Claire Kenyon. Efficient approximation and online algorithms: recent progress on classical combinatorical optimization problems and new applications, volume 3484 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32213-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32213-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3484.

Brim:2002:CCT

[BJKK02] Luboš Brim, Petr Jančar, Mojmír Křetínský, and Antonín Kuera, editors. CONCUR 2002

— Concurrency Theory: 13th International Conference, Brno, Czech Republic, August 20– 23, 2002: Proceedings, volume 2421 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44043-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2421.htm;
 http://www.springerlink.
com/content/978-3-540-
44043-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2421.

Babaoglu:2005:SSP

[BJM^{+}05] Özalp Babaoğlu, Márk Jelasity, Alberto Montresor, Christof Fetzer, Stefano Leonardi, Aad van Moorsel, and Maarten van Steen, editors. Self-star properties in complex information systems: conceptual and practical foundations, volume 3460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26009-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26009-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3460;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136551.

Berghammer:2009:RKA

[BJM09]  Rudolf  Berghammer,

Ali Mohamed Jaoua, and Bernhard Möller, editors. Relations and Kleene Algebra in Computer Science: 11th International Conference on Relational Methods in Computer Science, RelMiCS 2009, and 6th International Conference on Applications of Kleene Algebra, AKA 2009, Doha, Qatar, November 1–5, 2009. Proceedings, volume 5827 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04638-X (print), 3-642-04639-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04639-1.

Bliek:2003:GOC

[BJN03] Christian Bliek, Christophe Jermann, and Arnold Neumaier, editors. Global Optimization and Constraint Satisfaction: First International Workshop Global Constraint Optimization and Constraint Satisfaction, COCOS 2002, Valbonne-Sophia Antipolis, France, October 2–4, 2002: Revised Selected Papers, volume 2861 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20463-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C62 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2861.htm;
http://www.springerlink.
com/content/978-3-540-
20463-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2861;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94062.

Bredenfeld:2006:RRS

[BJNT06] Ansgar Bredenfeld, Adam Jacoff, Itsuki Noda, and Yasutake Takahashi, editors. RoboCup 2005:

Robot Soccer World Cup IX, volume 4020 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35437-9 (print), 3-540-35438-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35438-3.

Bjorner:1980:ASS

[Bjo80] D. Bjorner, editor. Abstract software specifications: 1979 Copenhagen Winter School, January 22–February 2, 1979: proceedings, volume 86 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10007-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.86. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0086.htm;
 http://www.springerlink.
com/content/978-0-387-
10007-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=86.

Sponsored by the Commission of the European Communities and others.

Bjorner:1987:VVF

[Bjo87] D. Bjorner, editor. VDM ’87: VDM, a formal method at work: VDM-Europe symposium 1987, Brussels, Belgium, March 23–26, 1987: proceedings, volume 252 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17654-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 V851 1987.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0252.
htm;   http://www.springerlink.
com/content/978-0-387-
17654-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=252.

Beldiceanu:2012:IAT

[BJP12] Nicolas Beldiceanu, Narendra Jussien, and Éric Pinson, editors. Integration of AI and OR Techniques in Contraint Programming for Combinatorial Optimzation Problems: 9th International Conference, CPAIOR 2012, Nantes, France, May 28 — June1, 2012. Proceedings, volume 7298 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29827-3 (print), 3-642-29828-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29828-8.

Butler:2006:RDC

[BJRT06] Michael Butler, Cliff B. Jones, Alexander Romanovsky, and Elena Troubitsyna, editors. Rigorous Development of Complex Fault-Tolerant Systems, volume 4157 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48265-2 (print), 3-540-48267-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        48267-3.

Butler:2009:MMT

[BJRT09] Michael Butler, Cliff Jones, Alexander Romanovsky, and Elena Troubitsyna, editors. Methods, Models and Tools for Fault Tolerance, volume 5454 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00866-6 (print), 3-642-00867-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00867-2.

Brewka:1993:NIL

[BJS93] Gerhard Brewka, K. P.

Janke, and P. H. Schmitt, editors. Nonmonotonic and inductive logic: second international workshop, Reinhardsbrunn Castle, Germany, December 2–6, 1991: proceedings, volume 659 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56433-0 (Berlin), 0-387-56433-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .N67 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0659.htm;
 http://www.springerlink.
com/content/978-3-540-
56433-1;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=659.

Proceedings of revised and extended papers presented at the Second International Workshop on Nonmonotonic and Inductive Logic, NIL ’91, which took place at Reinhardsbrunn Castle, Dec. 2–6, 1991.

Bohlen:1999:STD

[BJS99] Michael H. Böhlen, Christian S. Jensen, and Michel O. Scholl, editors. Spatio-temporal database management: International Workshop STDBM’99, Edinburgh, Scotland, September 10– 11, 1999: Proceedings, volume 1678 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66401-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1678; QA76.9.D3 S725 1999.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1678.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66401-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1678.

Bettini:2009:PLB

[BJSW09] Claudio Bettini, Sushil Jajodia, Pierangela Samarati, and X. Sean Wang, editors. Privacy in Location-Based Applications: Research Issues and Emerging Trends, volume 5599 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03510-8 (print), 3-642-03511-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03511-1.

Bibel:1980:CAD

[BK80] W. Bibel and Robert Kowalski, editors. 5th Conference on Automated Deduction, Les Arcs, France, July 8– 11, 1980, volume 87 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10009-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 .C66 1980. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0087.htm;

                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        10009-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=87.

Bolc:1981:DIP

[BK81] Leonard Bolc and Zenon Kulpa. Digital image processing systems, volume 109 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10705-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). iv + 353 pp. LCCN TA1632 .D495. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0109.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        10705-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=109.

Baeten:1990:CTC

[BK90] J. C. M. Baeten and J. W. Klop, editors. CONCUR ’90: theories of concurrency — unification and extension, Amsterdam, the Netherlands, August 27– 30, 1990: proceedings, volume 458 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 3-540-53048-7 (Berlin), 0-387-53048-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .C5927 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0458.htm;
 http://www.springerlink.
com/content/978-3-540-
53048-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=458.

“The ESPRIT Basic Research Action 3006, CONCUR (Theories of concurrency — unification and extension) started on September 1, 1989 … The conference CONCUR ’90, hosted by CWI with the help of the University of Amsterdam, marks the end of the first year of CONCUR”–Pref.

Borges:1996:AAI

[BK96] Dibio L. (Dibio Leander) Borges and Celso A. A. (Celso Antonio Alves) Kaestner, editors. Advances in artificial intelligence: 13th Brazilian Symposium on Artificial Intelligence, SBIA ’96, Curitiba, Brazil, October 23–25, 1996: proceedings, volume 1159 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-94900-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .S5434 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1159.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        94900-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1159.

Blostein:2002:GRA

[BK02] Dorothea Blostein and Young-Bin Kwon, editors. Graphics Recognition. Algorithms and Applications: 4th International Workshop, GREC 2001, Kingston, Ontario, Canada, September 7– 8, 2001. Selected Papers, volume 2390 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44066-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .I5538 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2390.htm;

http://www.springerlink.
com/content/978-3-540-
44066-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2390.

Bosch:2004:SRM

[BK04a] Jan Bosch and Charles Krueger, editors. Software Reuse: Methods, Techniques and Tools: 8th International Conference, ICSR 2004, Madrid, Spain, July 5– 9, 2004: Proceedings, volume 3107 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22335-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R47 I587 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3107.
htm;   http://www.springerlink.
com/content/978-3-540-
22335-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3107;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98465.

Brunner:2004:SMD

[BK04b] Marcus Brunner and Alexander Keller, editors. Self-Managing Distributed Systems: 14th IFIP/IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2003, Heidelberg, Germany, October 20–22, 2003: Proceedings, volume 2867 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20314-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I33835 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2867.
htm;   http://www.springerlink.
com/content/978-3-540-
20314-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2867;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94014.

Barley:2005:IAM

[BK05a] Michael Wayne Barley and Nikola K. Kasabov, editors. Intelligent agents and multi-agent systems: 7th Pacific Rim International Workshop on Multi-Agents, PRIMA 2004, Auckland, New Zealand, August 8–13, 2004: Revised selected papers, volume 3371 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25340-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 P75 2004. URL http:

//www.springerlink.
com/content/978-3-540-
25340-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3371;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107183.

Bierman:2005:DPL

[BK05b] Gavin Bierman and Christoph Koch, editors. Database programming languages: 10th International Symposium, DBPL 2005, Trondheim, Norway, August 28–29, 2005: revised selected papers, volume 3774 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30951-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5847 2005.

URL http://www.springerlink.
com/content/978-3-540-
30951-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3774.

Baras:2011:DGT

[BKA11] John S. Baras, Jonathan Katz, and Eitan Altman, editors. Decision and Game Theory for Security: Second International Conference, GameSec 2011, College Park, MD, Maryland, USA, November 14–15, 2011. Proceedings, volume 7037 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25279-6 (print), 3-642-25280-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25280-8.

Burakowski:2003:AQS

[BKB03] Wojciech Burakowski, Berthold Koch, and Andrzej Beben, editors. Architectures for quality of service in the Internet: International Workshop, Art-QoS 2003, Warsaw, Poland, March 24– 25, 2003: Revised Papers, volume 2698 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40444-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 A783 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2698.htm;
 http://www.springerlink.
com/content/978-3-540-
40444-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2698.

Bernstein:2009:SWI

[BKH^{+}09] Abraham Bernstein, David R.

Karger, Tom Heath, Lee Feigenbaum, Diana Maynard, Enrico Motta, and Krishnaprasad Thirunarayan, editors. The Semantic Web — ISWC 2009: 8th International Semantic Web Conference, ISWC 2009, Chantilly, VA, USA, October 25– 29, 2009. Proceedings, volume 5823 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04929-X (print), 3-642-04930-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04930-9.

Balandin:2011:SSN

[BKH11] Sergey Balandin, Yevgeni Koucheryavy, and Honglin Hu, editors. Smart Spaces and Next Generation Wired/Wireless Networking: 11th International Conference, NEW2AN 2011, and 4th Conference on Smart Spaces, ruSMART 2011, St. Petersburg, Russia, August 22–25, 2011. Proceedings, volume 6869 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22874-7 (print), 3-642-22875-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22875-9.

Bussler:2004:WIS

[BkHJ^{+}04] Christoph Bussler, Suk ki Hong, Woochun Jun, Roland Kaschek, Kinshuk, Shonali Krishnaswamy, Seng Wai Loke, Daniel Oberle, Debbie Richards, Amit Sharma, York Sure, and Bernhard Thalheim, editors. Web Information Systems— WISE 2004 Workshops: WISE 2004 International Workshops, Brisbane, Australia, November 22–24, 2004: Proceedings, volume 3307 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23892-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA168 .I57 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3307.htm;
 http://www.springerlink.
com/content/978-3-540-
23892-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3307;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b103295.

Bidoit:1991:ASS

[BKL^{+}91] M. Bidoit, H.-J. Kreowski, P. Lescanne, F. Orejas, and D. Sannella. Algebraic system specification and development: a survey and annotated bibliography, volume 501 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54060-1 (Berlin), 0-387-54060-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 98 pp. LCCN QA76.9.S88 A5 1991.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0501.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        54060-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=501.

Bestak:2012:NIIa

[BKL^{+}12a] Robert Bestak, Lukas Kencl, Li Erran Li, Joerg Widmer, and Hao Yin, editors. NETWORKING 2012: 11th International IFIP TC 6 Networking Conference, Prague, Czech Republic, May 21– 25, 2012, Proceedings, Part I, volume 7289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30044-8 (print), 3-642-30045-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30045-5.

Bestak:2012:NIIb

[BKL^{+}12b] Robert Bestak, Lukas Kencl, Li Erran Li, Joerg Widmer, and Hao Yin, editors. NETWORKING 2012: 11th International IFIP TC 6 Networking Conference, Prague, Czech Republic, May 21– 25, 2012, Proceedings, Part II, volume 7290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30053-7 (print), 3-642-30054-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30054-7.

Bouvry:2012:SII

[BKL^{+}12c] Pascal Bouvry, Mieczysław A.

Kłopotek, Franck Leprévost, Małgorzata Marciniak, Agnieszka Mykowiecka, and Henryk Rybiński, editors. Security and Intelligent Information Systems: International Joint Conferences, SIIS 2011, Warsaw, Poland, June 13–14, 2011, Revised Selected Papers, volume 7053 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25260-5 (print), 3-642-25261-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        25261-7.

Barkowsky:2007:SCV

[BKLM07] Thomas Barkowsky, Markus Knauff, Gérard Ligozat, and Daniel R. Montello, editors. Spatial Cognition V Reasoning, Action, Interaction: International Conference Spatial Cognition 2006, Bremen, Germany, September 24–28, 2006, Revised Selected Papers, volume 4387 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75665-5 (print), 3-540-75666-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75666-8.

Bohm:2011:ITB

[BKLP11] Christian Böhm, Sami Khuri, Lenka Lhotská, and Nadia Pisanti, editors. Information Technology in Bio- and Medical Informatics: Second International Conference, ITBAM 2011, Toulouse, France, August 31 — September 1, 2011. Proceedings, volume 6865 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23207-8 (print), 3-642-23208-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23208-4.

Bohm:2012:ITB

[BKLR12] Christian Böhm, Sami Khuri, Lenka Lhotská, and M. Elena Renda, editors. Information Technology in Bio- and Medical Informatics: Third International Conference, ITBAM 2012, Vienna, Austria, September 4–5, 2012. Proceedings, volume 7451 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32394-4 (print), 3-642-32395-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32395-9.

Boykov:2011:EMM

[BKLS11] Yuri Boykov, Fredrik Kahl, Victor Lempitsky, and Frank R. Schmidt, editors. Energy Minimization Methods in Computer Vision and Pattern Recognition: 8th International Conference, EMMCVPR 2011, St. Petersburg, Russia, July 25–27, 2011. Proceedings, volume 6819 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23093-8 (print), 3-642-23094-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23094-3.

Bouzeghoub:2001:NLP

[BKM01] M. Bouzeghoub, Zoubida Kedad, and Elisabeth Métais, editors. Natural language processing and information systems: 5th International Conference on Applications of Natural Language to Information Systems, NLDB 2000, Versailles, France, June 28–30, 2000: Revised Papers, volume 1959 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41943-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1959; QA76.9.N38 I558 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1959.htm;
http://www.springerlink.
com/content/978-3-540-
41943-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1959.

Broy:2006:ASC

[BKM06] Manfred Broy, Ingolf H.

Krüger, and Michael Meisinger, editors. Automotive Software — Connected Services in Mobile Networks: First Automotive Software Workshop, ASWSD 2004, San Diego, CA, USA, January 10–12, 2004, Revised Selected Papers, volume 4147 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37677-1 (print), 3-540-37678-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37678-1.

Brandstadt:2007:GTC

[BKM07] Andreas Brandstädt, Dieter Kratsch, and Haiko Müller, editors. Graph-Theoretic Concepts in Computer Science: 33rd International Workshop, WG 2007, Dornburg, Germany, June 21–23, 2007. Revised Papers, volume 4769 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74838-5 (print), 3-540-74839-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74839-7.

Bouguettaya:2008:SOC

[BKM08a] Athman Bouguettaya, Ingolf Krueger, and Tiziana Margaria, editors. Service-Oriented Computing — ICSOC 2008: 6th International Conference, Sydney, Australia, December 1–5, 2008. Proceedings, volume 5364 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89647-3 (print), 3-540-89652-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89652-4.

Broy:2008:MDD

[BKM08b] Manfred Broy, Ingolf H.

Krüger, and Michael Meisinger, editors. Model-Driven Development of Reliable Automotive Services: Second Automotive Software Workshop, ASWSD 2006, San Diego, CA, USA, March 15–17, 2006, Revised Selected Papers, volume 4922 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70929-0 (print), 3-540-70930-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70930-5.

Brusilovsky:2007:AWM

[BKN07] Peter Brusilovsky, Alfred Kobsa, and Wolfgang Nejdl, editors. The Adaptive Web: Methods and Strategies of Web Personalization, volume 4321 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72078-2 (print), 3-540-72079-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72079-9.

Breveglieri:2006:FDT

[BKNS06] Luca Breveglieri, Israel Koren, David Naccache, and Jean-Pierre Seifert, editors. Fault Diagnosis and Tolerance in Cryptography: Third International Workshop, FDTC 2006, Yokohama, Japan, October 10, 2006. Proceedings, volume 4236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46250-3 (print), 3-540-46251-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46251-4.

Butz:2003:SGT

[BKO03] Andreas Butz, Antonio Kruger, and Patrick Olivier, editors. Smart graphics: Third International Symposium, SG 2003, Heidelberg, Germany, July 2–4, 2003: Proceedings, volume 2733 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40557-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S378 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2733.htm;
 http://www.springerlink.
com/content/978-3-540-
40557-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2733.

Butz:2004:SGI

[BKO04] Andreas Butz, Antonio Krüger, and Patrick Olivier, editors. Smart Graphics: 4th International Symposium, SG 2004 Banff, Canada, May 23–25, 2004 Proceedings, volume 3031 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21977-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3031.htm;
 http://www.springerlink.
com/content/978-3-540-
21977-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3031;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b97744.

Benninghofen:1987:SR

[BKR87] B. (Benjamin) Benninghofen, S. Kemmerich, and Michael M. Richter. Systems of reductions, volume 277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18598-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 263 pp. LCCN QA267 .B461 1987. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0277.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        18598-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=277.

Benediktsson:2009:MCS

[BKR09] Jón Atli Benediktsson, Josef Kittler, and Fabio Roli, editors. Multiple Classifier Systems: 8th International Workshop, MCS 2009, Reykjavik, Iceland, June 10– 12, 2009. Proceedings, volume 5519 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02325-8 (print), 3-642-02326-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02326-2.

Brusilovsky:1996:MHV

[BKS96] Peter Brusilovsky, Piet A. M. Kommers, and N. A. (Norbert A.) Streitz. Multimedia, hypermedia, and virtual reality: models, systems, and applications, volume 1077 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61282-3. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 309 pp. LCCN QA76.575 .M38 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1077.htm;
http://www.springerlink.
com/content/978-3-540-
61282-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1077.

Banerjee:1999:HPC

[BKS99] Prithviraj Banerjee, V. K. Prasanna Kumar, and Bhabani P. Sinha, editors. High performance computing — HiPC’99: 6th International Conference, Calcutta, India, December 17–20, 1999: Proceedings, volume 1745 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66907-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I575 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1745.htm;
 http://www.springerlink.
com/content/978-3-540-
66907-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1745.

Bomarius:2001:PFS

[BKS01] Frank Bomarius and Seija Komi-Sirviö, editors. Product focused software process improvement: Third International Conference, PROFES 2001, Kaiserslautern, Germany, September 10– 13, 2001: Proceedings, volume 2188 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42571-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I5743 2001; QA267.A1 L43 no.2188. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2188.htm;
http://www.springerlink.
com/content/978-3-540-
42571-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2188.

Bomarius:2005:PFS

[BKS05] Frank Bomarius and Seija Komi-Sirviö, editors. Product Focused Software Process Improvement: 6th International Conference, PRO-FES 2005, Oulu, Finland, June 13–18, 2005. Proceedings, volume 3547 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26200-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .P7585 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26200-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3547.

Bordihn:2012:LAE

[BKT12] Henning Bordihn, Martin Kutrib, and Bianca Truthe, editors. Languages Alive: Essays Dedicated to Jürgen Dassow on the Occasion of His 65th Birthday, volume 7300 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31643-3 (print), 3-642-31644-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31644-9.

Blume:2010:FLP

[BKV10] Matthias Blume, Naoki Kobayashi, and Germán Vidal, editors. Functional and Logic Programming: 10th International Symposium, FLOPS 2010, Sendai, Japan, April 19–21, 2010. Proceedings, volume 6009 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12250-7 (print), 3-642-12251-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12251-4.

Bressan:2006:DES

[BKW06] Stéphane Bressan, Josef Küng, and Roland Wagner, editors. Database and Expert Systems Applications: 17th International Conference, DEXA 2006, Kraków, Poland, September 4– 8, 2006. Proceedings, volume 4080 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37871-5 (print), 3-540-37872-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37872-3.

Bhowmick:2008:DES

[BKW08] Sourav S. Bhowmick, Josef Küng, and Roland Wagner, editors. Database and Expert Systems Applications: 19th International Conference, DEXA 2008, Turin, Italy, September 1–5, 2008. Proceedings, volume 5181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85653-6 (print), 3-540-85654-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85654-2.

Bhowmick:2009:DES

[BKW09a] Sourav S. Bhowmick, Josef Küng, and Roland Wagner, editors. Database and Expert Systems Applications: 20th International Conference, DEXA 2009, Linz, Austria, August 31 — September 4, 2009. Proceedings, volume 5690 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03572-8 (print), 3-642-03573-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03573-9.

Bolc:2009:CVG

[BKW09b] Leonard Bolc, Juliusz L. Kulikowski, and Konrad Wojciechowski, editors. Computer Vision and Graphics: International Conference, ICCVG 2008 Warsaw, Poland, November 10– 12, 2008 Revised Papers, volume 5337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02344-4 (print), 3-642-02345-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02345-3.

Bravetti:2005:FTC

[BKZ05] Mario Bravetti, Leïla Kloul, and Gianluigi Zavattaro, editors. Formal Techniques for Computer Systems and Business Processes: European Performance Engineering Workshop, EPEW 2005 and International Workshop on Web Services and Formal Methods, WS-FM 2005, Versailles, France, September 1–3, 2005. Proceedings, volume 3670 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28701-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28701-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3670.

Boissonnat:1989:GRW

[BL89] J.-D. Boissonnat and J.- P. Laumond, editors. Geometry and robotics: workshop, Toulouse, France, May 26–28, 1988: proceedings, volume 391 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51683-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ210.3 .G46 1989. DM75.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0391.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51683-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=391.

“The workshop has been jointly organized by IN-RIA and LAAS-CNRS”– Pref.

Banatre:1992:RDH

[BL92] Jean-Pierre Banâtre and D. Le Métayer, editors. Research directions in high-level parallel programming languages: Mont Saint-Michel, France, June 17–19, 1991: proceedings, volume 574 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55160-3 (Berlin), 0-387-55160-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .R47 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0574.htm;
 http://www.springerlink.
com/content/978-3-540-
55160-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=574.

Proceedings of a workshop held at Mont Saint-Michel.

Banatre:1994:HSA

[BL94] Michel Banâtre and P. A.

Lee, editors. Hardware and software architectures for fault tolerance: experiences and perspectives, volume 774 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57767-X (Berlin), 0-387-57767-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F38 H36 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0774.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        57767-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=774.

Proceedings of a workshop held at Le Mont Saint Michel, France in June, 1993.

Brandstadt:2001:GTC

[BL01] Andreas Brandstädt and Van Bang Le, editors. Graph-theoretic concepts in computer science: 27th International Workshop, WG 2001, Boltenhagen, Germany, June 14–16, 2001: Proceedings, volume 2204 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42707-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G715 2001; QA267.A1 L43 no.2204. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2204.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42707-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2204.

Bosnacki:2002:MCS

[BL02a] Dragan Bosnacki and Stefan Leue, editors.

Model checking software: 9th International SPIN Workshop, Grenoble, France, April 11– 13, 2002: Proceedings, volume 2318 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43477-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 S65 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2318.htm;
 http://www.springerlink.
com/content/978-3-540-
43477-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2318.

Brinksma:2002:CAV

[BL02b] Ed Brinksma and K. Guldstrand Larsen, editors. Computer Aided Verification: 14th International Conference, CAV 2002, Copenhagen, Denmark, July 27–31, 2002. Proceedings, volume 2404 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43997-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E93 C38 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2404.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43997-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2404.

Blundo:2003:TCS

[BL03] Carlo Blundo and Cosimo Laneve, editors. Theoretical Computer Science: 8th Italian Conference, ICTCS 2003, Bertinoro, Italy, October 13–15, 2003: Proceedings, volume 2841 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20216-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I185 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2841.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20216-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2841;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13810.

Bazzan:2004:AAI

[BL04a] Ana L. C. Bazzan and Sofiane Labidi, editors. Advances in artificial intelligence: SBIA 2004, 17th Brazilian Symposium on Artificial Intelligence, São Luis, Maranhão, Brazil, September 29–October 1, 2004: proceedings, volume 3171 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23237-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .S5434 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23237-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3171;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100195.

Bruynooghe:2004:PDC

[BL04b] Maurice Bruynooghe and Kung-Kiu Lau, editors. Program Development in Computational Logic: A Decade of Research Advances in Logic-Based Program Development, volume 3049 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22152-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .P74 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3049.htm;
 http://www.springerlink.
com/content/978-3-540-
22152-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3049;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98187.

Baird:2005:HIP

[BL05a] Henry S. Baird and Daniel P. Lopresti, editors. Human interactive proofs: first international workshop, HIP 2005, Bethlehem, PA, USA, May 19–20, 2005, proceedings, volume 3517 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26001-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-

26001-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3517;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136509.

Beigl:2005:SAO

[BL05b] Michael Beigl and Paul Lukowicz, editors. Systems aspects in organic and pervasive computing—ARCS 2005: 18th International Conference on Architecture of Computing Systems, Innsbruck, Austria, March 14–17, 2005: Proceedings, volume 3432 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25273-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A73 A736 2005.

URL http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3432;
 http://www.springerlink.
com/content/978-3-540-
25273-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3432;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106632.

Brodal:2005:AEA

[BL05c] Gerth Stølting Brodal and Stefano Leonardi, editors. Algorithms — ESA 2005: 13th annual European symposium, Palma de Mallorca, Spain, October 3–6, 2005, proceedings, volume 3669 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29118-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29118-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3669.

Barua:2006:PCI

[BL06a] Rana Barua and Tanja Lange, editors. Progress in Cryptology — IN-DOCRYPT 2006: 7th International Conference on Cryptology in India, Kolkata, India, December 11–13, 2006. Proceedings, volume 4329 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49767-6 (print), 3-540-49769-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49769-1.

Bodlaender:2006:PEC

[BL06b] Hans L. Bodlaender and Michael A. Langston, editors. Parameterized and Exact Computation: Second International Workshop, IWPEC 2006, Zürich, Switzerland, September 13–15, 2006. Proceedings, volume 4169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39098-7 (print), 3-540-39101-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
39101-2.

Bres:2006:VII

[BL06c] Stéphane Bres and Robert Laurini, editors. Visual information and information systems: 8th international conference, VISUAL 2005, Amsterdam, The Netherlands, July 5, 2005: revised selected papers, volume 3736 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30488-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.I6 V558 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        30488-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3736.

Buschkes:2006:DIM

[BL06d] Roland Büschkes and Pavel Laskov, editors. Detection of Intrusions and Malware & Vulnerability Assessment: Third International Conference, DIMVA 2006, Berlin, Germany, July 13–14, 2006. Proceedings, volume 4064 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36014-X (print), 3-540-36017-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        36017-9.

Biskup:2007:CSE

[BL07a] Joachim Biskup and Javier López, editors. Computer Security — ESORICS 2007: 12th European Symposium On Research In Computer Security, Dresden, Germany, September 24 — 26, 2007. Proceedings, volume 4734 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74834-2 (print), 3-540-74835-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74835-9.

Boztas:2007:AAA

[BL07b] Serdar Boztaş and Hsiao-Feng (Francis) Lu, editors. Applied Algebra, Algebraic Algorithms and Error-Correcting Codes: 17th International Symposium, AAECC-17, Bangalore, India, December 16–20, 2007. Proceedings, volume 4851 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77223-5 (print), 3-540-77224-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        77224-8.

Bellahsene:2008:AIS

[BL08] Zohra Bellahsène and Michel Léonard, editors. Advanced Information Systems Engineering: 20th International Conference, CAiSE 2008 Montpellier, France, June 16– 20, 2008 Proceedings, volume 5074 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69533-8 (print), 3-540-69534-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69534-9.

Bernstein:2010:PCA

[BL10a] Daniel J. Bernstein and Tanja Lange, editors. Progress in Cryptology — AFRICACRYPT 2010: Third International Conference on Cryptology in Africa, Stellenbosch, South Africa, May 3– 6, 2010. Proceedings, volume 6055 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12677-4 (print), 3-642-12678-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12678-9.

Bosch:2010:SPL

[BL10b] Jan Bosch and Jaejoon Lee, editors. Software Product Lines: Going Beyond: 14th International Conference, SPLC 2010, Jeju Island, South Korea, September 13– 17, 2010. Proceedings, volume 6287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15578-2 (print), 3-642-15579-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15579-6.

Butz:2011:AAI

[BL11] Cory Butz and Pawan Lingras, editors. Advances in Artificial Intelligence: 24th Canadian Conference on Artificial Intelligence, Canadian AI 2011, St. John’s, Canada, May 25–27, 2011. Proceedings, volume 6657 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21042-2 (print), 3-642-21043-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21043-3.

Barbosa:2012:FAC

[BL12] Luís Soares Barbosa and Markus Lumpe, editors. Formal Aspects of Component Software: 7th International Workshop, FACS 2010, Guimarães, Portugal, October 14–16, 2010, Revised Selected Papers, volume 6921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27268-1 (print), 3-642-27269-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27269-1.

Blaser:1980:DBT

[Bla80] A. Blaser, editor. Data base techniques for pictorial applications, Florence, June 20–22, 1979, volume 81 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CODEN LNCSD9. ISBN 0-387-09763-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .D36. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0081.
htm;   http://www.springerlink.
com/content/978-0-387-
09763-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=81.

Proceedings of a conference sponsored by IBM Europe and IBM Italy.

Blaser:1988:NLC

[Bla88] A. Blaser, editor. Natural language at the computer: proceedings, volume 320 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 3-540-50011-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 S381 1988.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0320.
htm;   http://www.springerlink.
com/content/978-3-540-
50011-7;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=320.

Contents: Introduction / A. Blaser — Computational linguistics: issues and solutions / M. King — From feature bundles to abstract data types: new directions in the representation and processing of linguistic knowledge / H. Uszkoreit — Issues in parsing / K. Jensen — Discourse representation theory: what it is and where it ought to go / H. Kamp — The LEX project-concepts and results / H. Lehmann — From sentences to discourse: some aspects of the computational treatment of language / F. Guenthner — Computational linguistics at the Heidelberg Science Center / M. Zoeppritz.

Blaser:1990:DSI

[Bla90] A. Blaser, editor. Database systems of the 90s: International Symposium, Muggelsee, Berlin, FRG, November 5–7, 1990: proceedings, volume 466 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-53397-4 (Berlin), 0-387-53397-4 (New York).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D3589 1990. DM47.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0466.
htm;   http://www.springerlink.
com/content/978-3-540-
53397-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=466.

Symposium sponsored by IBM Germany.

Blackburn:2003:MUC

[Bla03a] Patrick Blackburn, editor. Modeling and using context: 4th International and Interdisciplinary Conference, CONTEXT 2003, Stanford, CA, USA, June 23–25, 2003: Proceedings, volume 2680 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40380-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65 I537 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2680.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        40380-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2680.

Blaze:2003:FCI

[Bla03b] Matt Blaze, editor. Financial cryptography: 6th International Conference, FC 2002, Southampton, Bermuda, March 11–14, 2002: Revised Papers, volume 2357 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00646-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2357.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00646-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2357.

Black:2005:EOO

[Bla05] Andrew P. Black, editor. ECOOP 2005 — objectoriented programming: 19th European conference, Glasgow, UK, July 25–29, 2005: proceedings, volume 3586 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27992-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 2005. URL http:

//www.springerlink.
com/content/978-3-540-
27992-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3586.

Berthold:2003:AID

[BLB^{+}03] Michael R. Berthold, Hans-Joachim Lenz, Elizabeth Bradley, Rudolf Kruse, and Christian Borgelt, editors. Advances in Intelligent Data Analysis V: 5th International Conference on Intelligent Data Analysis, IDA 2003, Berlin, Germany, August 28– 30, 2003: Proceedings, volume 2810 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40813-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA276.A1. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2810.htm;
 http://www.springerlink.
com/content/978-3-540-
40813-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=2810;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13240.

Blikle:1975:MFC

[Bli75] Andrzej Blikle, editor. Mathematical foundations of computer science: 3rd symposium at Jadwisin near Warsaw, June 17–22, 1974, volume 28 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07162-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.S4 S989 1974.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0028.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        07162-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=28.

“Organized by the Computation Center of the Polish Academy of Sciences and sponsored by the International Mathematical Stefan Banach Center in Warsaw.”.

Blikle:1987:MPT

[Bli87] Andrzej Blikle. Meta-Soft primer: towards a metalanguage for applied denotational semantics, volume 288 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18657-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 140 pp. LCCN QA76.7 .B571 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0288.htm;
 http://www.springerlink.
com/content/978-0-387-
18657-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=288.

Bentley:2008:AIS

[BLJ08] Peter J. Bentley, Doheon Lee, and Sungwon Jung, editors. Artificial Immune Systems: 7th International Conference, ICARIS 2008, Phuket, Thailand, August 10–13, 2008. Proceedings, volume 5132 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85071-6 (print), 3-540-85072-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        85072-4.

Bergmann:2008:MST

[BLKP08] Ralph Bergmann, Gabriela Lindemann, Stefan Kirn, and Michal Pěchouček, editors. Multiagent System Technologies: 6th German Conference, MATES 2008, Kaiserslautern, Germany, September 23– 26, 2008. Proceedings, volume 5244 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87804-1 (print), 3-540-87805-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87805-6.

Bode:2001:EPP

[BLKW01] Arndt Bode, Thomas Ludwig, Wolfgang Karl, and Roland Wismüller, editors. Euro-Par 2000 parallel processing: 6th International Euro-Par Conference, Munich, Germany, August 1– September 1, 2000: Proceedings, volume 1900 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-67956-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E865 2000; QA76.58 .I553 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1900.htm;
http://www.springerlink.
com/content/978-3-540-
67956-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1900.

Baltes:2010:RRS

[BLNG10] Jacky Baltes, Michail G.

Lagoudakis, Tadashi Naruse, and Saeed Shiry Ghidary, editors. RoboCup 2009: Robot Soccer World Cup XIII, volume 5949 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11875-5 (print), 3-642-11876-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11876-0.

Bao:2007:CNS

[BLO^{+}07] Feng Bao, San Ling, Tatsuaki Okamoto, Huaxiong Wang, and Chaoping Xing, editors. Cryptology and Network Security: 6th International Conference, CANS 2007, Singapore, December 8– 10, 2007. Proceedings, volume 4856 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76968-4 (print), 3-540-76969-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76969-9.

Bulthoff:2002:BMC

[BLPW02] Heinrich H. Bülthoff, S.- W. Lee, T. A. Poggio, and C. Wallraven, editors. Biologically Motivated Computer Vision: Second International Workshop, BMCV 2002, Tübingen, Germany, November 22– 24, 2002. Proceedings, volume 2525 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00174-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP474 .I446 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2525.htm;

 http://www.springerlink.
com/content/978-3-540-
00174-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2525.

Also available via the World Wide Web.

Baeten:2003:ALP

[BLPW03] Jos C. M. Baeten, Jan Karel Lenstra, Joachim Parrow, and Gerhard J. Woeginger, editors. Automata, languages and programming: 30th international colloquium, ICALP 2003, Eindhoven, The Netherlands, June 30–July 4, 2003: Proceedings, volume 2719 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40493-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I551 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2719.htm;
 http://www.springerlink.
com/content/978-3-540-
40493-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2719.

30th International Colloquium on Automata, Languages and Programming (ICALP 2003).

Bustard:2002:SWC

[BLS02] David Bustard, Weiru Liu, and Roy Sterritt, editors. Soft-ware 2002: computing in an imperfect world: First International Conference, Soft-Ware 2002, Belfast, Northern Ireland, April 8–10, 2002: Proceedings, volume 2311 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43481-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .S616 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2311.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43481-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2311.

Balcazar:2006:ALT

[BLS06] José L. Balcázar, Philip M. Long, and Frank Stephan, editors. Algorithmic Learning Theory: 17th International Conference, ALT 2006, Barcelona, Spain, October 7–10, 2006. Proceedings, volume 4264 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46649-5 (print), 3-540-46650-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46650-5.

Barthe:2012:ESS

[BLS12] Gilles Barthe, Benjamin Livshits, and Riccardo Scandariato, editors. Engineering Secure Software and Systems: 4th International Symposium, ESSoS 2012, Eindhoven, The Netherlands, February, 16–17, 2012. Proceedings, volume 7159 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28165-6 (print), 3-642-28166-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28166-2.

Bezhanishvili:2011:LLC

[BLSS11] Nick Bezhanishvili, Sebastian Löbner, Kerstin Schwabe, and Luca Spada, editors. Logic, Language, and Computation: 8th International Tbilisi Symposium on Logic, Language, and Computation, TbiLLC 2009, Bakuriani, Georgia, September 21–25, 2009. Revised Selected Papers, volume 6618 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22302-8 (print), 3-642-22303-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22303-7.

Braquelaire:2002:DGC

[BLV02] Achille Braquelaire, Jacques-Olivier Lachaud, and Anne Vialard, editors. Discrete geometry for computer imagery: 10th International Conference, DGCI 2002, Bordeaux, France, April 3– 5, 2002: Proceedings, volume 2301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43380-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .D525 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2301.

htm;   http://www.springerlink.
com/content/978-3-540-
43380-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2301.

Bonanno:2010:LFG

[BLvdH10] Giacomo Bonanno, Benedikt Löwe, and Wiebe van der Hoek, editors. Logic and the Foundations of Game and Decision Theory — LOFT 8: 8th International Conference, Amsterdam, The Netherlands, July 3–5, 2008, Revised Selected Papers, volume 6006 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15163-9 (print), 3-642-15164-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15164-4.

Burkhard:2007:MAS

[BLVV07] Hans-Dieter Burkhard, Gabriela Lindemann, Rineke Verbrugge, and László Zsolt Varga, editors. Multi-Agent Systems and Applications V: 5th International Central and Eastern European Conference on Multi-Agent Systems, CEEMAS 2007, Leipzig, Germany, September 25– 27, 2007. Proceedings, volume 4696 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75253-6 (print), 3-540-75254-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75254-7.

Branquart:1982:ADC

[BLW82] Paul Branquart, Georges Louis, and Pierre Wodon. An analytical description of CHILL, the CCITT high level language, volume 128 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11196-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 277 pp. LCCN QA76.73.C18 B7 1982. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0128.htm;
 http://www.springerlink.
com/content/978-0-387-
11196-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=128.

Burns:1987:RAT

[BLW87] Alan Burns, A. Lister, and Andrew J. Wellings. A review of Ada tasking, volume 262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-18008-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 141 pp. LCCN QA76.73.A35 B871 1987.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0262.
htm;   http://www.springerlink.
com/content/978-0-387-
18008-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=262.

Bao:2009:ISP

[BLW09] Feng Bao, Hui Li, and Guilin Wang, editors. Information Security Practice and Experience: 5th International Conference, ISPEC 2009 Xi’an, China, April 13–15, 2009 Proceedings, volume 5451 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00842-9 (print), 3-642-00843-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00843-6.

Bouchitte:1994:OAA

[BM94] Vincent Bouchitte and Michel Morvan, editors. Orders, algorithms, and applications: International Workshop ORDAL ’94, Lyon, France, July 1994: proceedings, volume 831 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58274-6, 0-387-58274-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA171.48 .I56 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0831.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58274-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=831.

Summary in French.

Babaoglu:1996:DAI

[BM96] Ozalp Babaoglu and Keith Marzullo, editors. Distributed algorithms: 10th International Workshop, WDAG ’96, Bologna, Italy, October 9–11, 1996: proceedings, volume 1151 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61769-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I563 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1151.
htm;   http://www.springerlink.
com/content/978-3-540-
61769-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1151.

Bosch:1998:OOT

[BM98] Jan Bosch and Stuart Mitchell, editors. Object-oriented technology: ECOOP ’97 workshop reader, ECOOP ’97 workshops, Jyväskylä, Finland, June 9–13, 1997: proceedings, volume 1357 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64039-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35 O26 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1357.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        64039-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1357.

Baeten:1999:CCT

[BM99a] J. C. M. Baeten and S. (Sjouke) Mauw, editors. CONCUR’99: concurrency theory: 10th International Conference, Eindhoven, The Netherlands, August 24– 27, 1999 proceedings, volume 1664 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66425-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1664; QA76.58 .I53 1999. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1664.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66425-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1664.

Bezivin:1999:UML

[BM99b] Jean Bézivin and Pierre-Alain Muller, editors. The unified modeling language: UML ’98: beyond the notation: First International Workshop, Mulhouse, France, June 3–4, 1998: Selected Papers, volume 1618 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66252-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35 U54 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1618.
htm;   http://www.springerlink.
com/content/978-3-540-
66252-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1618.

Boiten:2002:MPC

[BM02a] Eerke Boiten and Bernhard Möller, editors. Mathematics of Program Construction: 6th International Conference, MPC 2002, Dagstuhl Castle, Germany, July 8–10, 2002. Proceedings, volume 2386 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43857-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .M3661 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2386.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        43857-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2386.

Bose:2002:ACI

[BM02b] Prosenjit Bose and Pat Morin, editors. Algorithms and Computation: 13th International Symposium, ISAAC 2002, Vancouver, BC, Canada, November 21–23, 2002. Proceedings, volume 2518 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00142-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I73 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2518.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        00142-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2518.

Also available via the

World Wide Web.

Baaz:2003:CSL

[BM03a] Matthias Baaz and Johann A. Makowsky, editors. Computer Science Logic: 17th International Workshop, CSL 2003, 12th Annual Conference of the EACSL, and the 8th Kurt Goedel Colloquium, KGC 2003, Vienna, Austria, August 25–30, 2003: Proceedings, volume 2803 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40801-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.L63 C68 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2803.
htm;   http://www.springerlink.
com/content/978-3-540-
40801-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2803;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13224.

Bidoit:2003:CUM

[BM03b] Michel Bidoit and Peter D. Mosses, editors. CASL User Manual: Introduction to using the Common Algebraic Specification Language CASL, volume 2900 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20766-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.C158. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2900.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20766-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2900;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11968.

Boyd:2003:ISI

[BM03c] Colin Boyd and Wenbo Mao, editors. Information Security: 6th International Conference, ISC 2003, Bristol, UK, October 1–3, 2003: Proceedings, volume 2851 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20176-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I85 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2851.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20176-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2851;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=

volume&id=doi:10.1007/
b13477.

Bandini:2005:AAA

[BM05a] Stefania Bandini and Sara Manzoni, editors. AI*IA 2005: Advances in Artificial Intelligence: 9th Congress of the Italian Association for Artificial Intelligence, Milan, Italy, September 21–23, 2005. Proceedings, volume 3673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29041-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29041-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3673.

Bartak:2005:IAT

[BM05b] Roman Barták and Michela Milano, editors. Integration of AI and OR techniques in constraint programming for combinatorial problems: second international conference, CPAIOR 2005, Prague Czech Republic, May 31–June 1, 2005. Proceedings, volume 3524 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26152-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q340 .I56 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        26152-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3524;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136920.

Bucher:2006:ABI

[BM06] Philipp Bücher and Bernard M. E. Moret, editors. Algorithms in Bioinformatics: 6th International Workshop, WABI 2006, Zurich, Switzerland, September 11–13, 2006. Proceedings, volume 4175 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39583-0 (print), 3-540-39584-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        39584-3.

Bouajjani:2009:CAV

[BM09a] Ahmed Bouajjani and Oded Maler, editors. Computer Aided Verification: 21st International Conference, CAV 2009, Grenoble, France, June 26 — July 2, 2009. Proceedings, volume 5643 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02657-5 (print), 3-642-02658-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02658-4.

Bry:2009:STW

[BM09b] François Bry and Jan Małuszyński, editors. Semantic Techniques for the Web: The REW-ERSE Perspective, volume 5500 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04580-4 (print), 3-642-04581-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04581-3.

Bichindaritz:2010:CBR

[BM10] Isabelle Bichindaritz and Stefania Montani, editors. Case-Based Reasoning. Research and Development: 18th International Conference on Case-Based Reasoning, ICCBR 2010, Alessandria, Italy, July 19– 22, 2010. Proceedings, volume 6176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14273-7 (print), 3-642-14274-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14274-1.

Beckert:2011:FVO

[BM11a] Bernhard Beckert and Claude Marché, editors. Formal Verification of Object-Oriented Software: International Conference, FoVeOOS 2010, Paris, France, June 28–30, 2010, Revised Selected Papers, volume 6528 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18069-8 (print), 3-642-18070-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18070-5.

Brunnler:2011:ARA

[BM11b] Kai Brünnler and George Metcalfe, editors. Automated Reasoning with Analytic Tableaux and Related Methods: 20th International Conference, TABLEAUX 2011, Bern, Switzerland, July 4–8, 2011. Proceedings, volume 6793 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22118-1 (print), 3-642-22119-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22119-4.

Buchanan:2008:DLU

[BMC08] George Buchanan, Masood Masoodian, and Sally Jo Cunningham, editors. Digital Libraries: Universal and Ubiquitous Access to Information: 11th International Conference on Asian Digital Libraries, ICADL 2008, Bali, Indonesia, December 2–5, 2008. Proceedings, volume 5362 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89532-9 (print), 3-540-89533-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89533-6.

Bouchou-Markhoff:2011:IAA

[BMCCM11] Béatrice Bouchou-Markhoff, Pascal Caron, Jean-Marc Champarnaud, and Denis Maurel, editors. Implementation and Application of Automata: 16th International Conference, CIAA 2011, Blois, France, July 13– 16, 2011. Proceedings, volume 6807 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22255-2 (print), 3-642-22256-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22256-6.

Baumeister:2005:EPA

[BMH05] Hubert Baumeister, Michele Marchesi, and Mike Holcombe, editors. Extreme programming and agile processes in software engineering: 6th international conference, XP 2005, Sheffield, UK, June 18–23, 2005, proceedings, volume 3556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26277-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 X62 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26277-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3556;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b137278.

Bloomfield:1988:VVW

[BMJ88] R. Bloomfield, L. Marshall, and R. Jones, editors. VDM ’88: VDM, the way ahead: 2nd VDM-Europe symposium, Dublin, Ireland, September, 1988: proceedings, volume 328 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 3-540-50214-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 V8541 1988.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0328.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        50214-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=328.

Balandin:2008:NGT

[BMK08] Sergey Balandin, Dmitri Moltchanov, and Yevgeni Koucheryavy, editors. Next Generation Teletraffic and Wired/ Wireless Advanced Networking: 8th International Conference, NEW2AN and 1st Russian Conference on Smart Spaces, ruSMART 2008 St. Petersburg, Russia, September 3–5, 2008. Proceedings, volume 5174 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85499-1 (print), 3-540-85500-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85500-2.

Balandin:2009:SSN

[BMK09] Sergey Balandin, Dmitri Moltchanov, and Yevgeni Koucheryavy, editors. Smart Spaces and Next Generation Wired/ Wireless Networking: 9th International Conference, NEW2AN 2009 and Second Conference on Smart Spaces, ruS-MART 2009, St. Petersburg, Russia, September 15–18, 2009. Proceedings, volume 5764 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04188-4 (print), 3-642-04190-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04190-7.

Boavida:2007:WWI

[BMM07] Fernando Boavida, Edmundo Monteiro, and Saverio Mascolo, editors. Wired/wireless Internet communications: 5th international conference, WWIC 2007, Coimbra, Portugal, May 23–25, 2007. proceedings, volume 4517 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-72694-2 (softcover). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        72694-4.

Bolc:2005:IMT

[BMN05] Leonard Bolc, Zbigniew Michalewicz, and Toyoaki Nishida, editors. Intelligent Media Technology for Communicative Intelligence: Second International Workshop, IMTCI 2004, Warsaw, Poland, September 13–14, 2004. Revised Selected Papers, volume 3490 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29035-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29035-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3490.

Boavida:2002:PSI

[BMO02] Fernando Boavida, Edmundo Monteiro, and Joao Orvalho, editors. Protocols and Systems for Interactive Distributed Multimedia: Joint International Workshops on Interactive Distributed Multimedia Systems and Protocols for Multimedia Systems, IDMS/PROMS 2002, Coimbra, Portugal, November 26–29, 2002. Proceedings, volume 2515 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00169-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I59 I586 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2515.
htm;   http://www.springerlink.
com/content/978-3-540-
00169-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2515.

Bergman:1990:ALU

[BMP90] C. H. Bergman, R. D.

Maddux, and Don Pigozzi, editors. Algebraic logic and universal algebra in computer science: conference, Ames, Iowa, USA, June 1–4, 1988: proceedings, volume 425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-97288-9 (New York), 3-540-97288-9 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.425. DM47.00. URL

http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t0425.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        97288-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=425.

Butscher:1996:IDM

[BMP96] Berthold Butscher, Eckhard Moeller, and Herwart Pusch, editors. Interactive distributed multimedia systems and services: European Workshop IDMS ’96, Berlin, Germany, March 4– 6, 1996: proceedings, volume 1045 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60938-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I57 E97 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1045.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60938-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1045.

Bauknecht:2000:ECW

[BMP00] Kurt Bauknecht, Sanjay Kumar Madria, and G. Pernul, editors. Electronic Commerce and Web Technologies: First International Conference, EC-Web 2000, London, UK, September 4–6, 2000. Proceedings, volume 1875 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67981-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.32 .E35 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1875.htm;
 http://www.springerlink.
com/content/978-3-540-
67981-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1875.

Bauknecht:2001:ECW

[BMP01] Kurt Bauknecht, Sanjay Kumar Madria, and G. Pernul, editors. Electronic commerce and Web technologies: Second International Conference, EC-WEB 2001, Munich, Germany, September 4–6, 2001: Proceedings, volume 2115 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42517-9 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2115; HF5548.32 .E35 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2115.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42517-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2115.

The refereed proceedings of the Second International Conference on Electronic Commerce and Web Technologies.

Barnaghi:2009:SSC

[BMPM09] Payam Barnaghi, Klaus Moessner, Mirko Presser, and Stefan Meissner, editors. Smart Sensing and Context: 4th European Conference, EuroSSC 2009, Guildford, UK, September 16–18, 2009. Proceedings, volume 5741 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04470-0 (print), 3-642-04471-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04471-7.

Bellahsene:2004:DXT

[BMR^{+}04] Zohra Bellahsène, Tova Milo, Michael Rys, Dan Suciu, and Rainer Unland, editors. Database and XML Technologies: Second International XML Database Symposium, XSYM 2004, Toronto, Canada, August 29–30, 2004: Proceedings, volume 3186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22969-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3186.htm;
http://www.springerlink.
com/content/978-3-540-
22969-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3186;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100256.

Birman:1995:TPD

[BMS95] Kenneth P. Birman, Friedemann Mattern, and Andre Schiper, editors. Theory and practice in distributed systems: International Workshop, Dagstuhl Castle, Germany, September 5– 9, 1994: Selected Papers, volume 938 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60042-6 (Berlin), 0-387-60042-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.938. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0938.htm;
 http://www.springerlink.
com/content/978-3-540-
60042-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=938.

Broy:1996:FSS

[BMS96] M. Broy, Stephan Merz, and Katharina Spies. Formal systems specification: the RPC-memory specification case study, volume 1169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61984-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xxiii + 540 pp. LCCN TK7895.M4 F67 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1169.htm;

http://www.springerlink.
com/content/978-3-540-
61984-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1169.

Brewster:2001:HHC

[BMS01] Stephen Brewster and Roderick Murray-Smith, editors. Haptic humancomputer interaction: First International Workshop, Glasgow, UK, August 31–September 1, 2000: Proceedings, volume 2058 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42356-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 H37 2000; QA267.A1 L43 no.2058. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2058.htm;
http://www.springerlink.
com/content/978-3-540-
42356-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2058.

Berghammer:2004:RKA

[BMS04a]  Rudolf  Berghammer,

Bernhard Möller, and Georg Struth, editors. Relational and Kleene-Algebraic Methods in Computer Science: 7th International Seminar on Relational Methods in Computer Science and 2nd International Workshop on Applications of Kleene Algebra, Bad Malente, Germany, May 12–17, 2003: revised Selected Papers, volume 3051 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22145-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA10 .I58 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3051.htm;
 http://www.springerlink.
com/content/978-3-540-
22145-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3051;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98117.

Blackwell:2004:DRI

[BMS04b] Alan Blackwell, Kim Marriott, and Atsushi Shimojima, editors. Diagrammatic Representation and Inference: Third International Conference, Diagrams 2004 Cambridge, UK, March 22–24, 2004: Proceedings, volume 2980 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21268-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA90 .D53 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2980.htm;
 http://www.springerlink.
com/content/978-3-540-
21268-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2980;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95854.

Bailey:2008:WIS

[BMS^{+}08a] James Bailey, David Maier, Klaus-Dieter Schewe, Bernhard Thalheim, and Xiaoyang Sean Wang, editors. Web Information Systems Engineering

— WISE 2008: 9th International Conference, Auckland, New Zealand, September 1–3, 2008. Proceedings, volume 5175 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85480-0 (print), 3-540-85481-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85481-4.

Berghammer:2008:RKA

[BMS08b]  Rudolf  Berghammer,

Bernhard Möller, and Georg Struth, editors. Relations and Kleene Algebra in Computer Science: 10th International Conference on Relational Methods in Computer Science, and 5th International Conference on Applications of Kleene Algebra, RelMiCS/AKA 2008, Frauenwörth, Germany, April 7–11, 2008. Proceedings, volume 4988 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78912-X (print), 3-540-78913-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78913-0.

Berekovic:2009:ACS

[BMSHW09] Mladen Berekovic, Christian Müller-Schloer, Christian Hochberger, and Stephan Wong, editors. Architecture of Computing Systems — ARCS 2009: 22nd International Conference, Delft, The Netherlands, March 10– 13, 2009. Proceedings, volume 5455 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00453-9 (print), 3-642-00454-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00454-4.

Barreiro:2004:BMD

[BMSMS04] José M. Barreiro, Fernando Martin-Sanchez, Víctor Maojo, and Ferran Sanz, editors. Biological and medical data analysis: 5th international symposium, ISN-MDA 2004, Barcelona, Spain, November 18– 19, 2004, proceedings, volume 3337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23964-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R853.S7 I582 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23964-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3337;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104033.

Bleris:2012:BRA

[BMSW12] Leonidas Bleris, Ion Măndoiu, Russell Schwartz, and Jianxin Wang, editors. Bioinformatics Research and Applications: 8th International Symposium, ISBRA 2012, Dallas, TX, USA, May 21–23, 2012. Proceedings, volume 7292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30190-8 (print), 3-642-30191-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30191-9.

Barley:2009:SSM

[BMU^{+}09] Mike Barley, Haralambos Mouratidis, Amy Unruh, Diana Spears, Paul Scerri, and Fabio Massacci, editors. Safety and Security in Multiagent Systems: Research Results from 2004-2006, volume 4324 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04878-1 (print), 3-642-04879-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04879-1.

Bandini:2010:CAI

[BMUV10] Stefania Bandini, Sara Manzoni, Hiroshi Umeo, and Giuseppe Vizzari, editors. Cellular Automata: 9th International Conference on Cellular Automata for Research and Industry, ACRI 2010, Ascoli Piceno, Italy, September 21–24, 2010. Proceedings, volume 6350 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15978-8 (print), 3-642-15979-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15979-4.

Bird:1993:MPC

[BMW93] R. S. Bird, Carroll Morgan, and Jim Woodcock, editors. Mathematics of program construction: second international conference, Oxford, UK, June 29–July 3, 1992: proceedings, volume 669 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56625-2 (Berlin), 0-387-56625-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .M367 1993. DM72.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0669.
htm;   http://www.springerlink.
com/content/978-3-540-
56625-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=669.

“Papers … presented at the Second International Conference on the Mathematics of Program Construction, held at St Catherine’s College, Oxford, during the week of 29 June-3 July, 1992”– Pref.

Bolc:2010:SIS

[BMW10] Leonard Bolc, Marek Makowski, and Adam Wierzbicki, editors. Social Informatics: Second International Conference, SocInfo 2010, Laxenburg, Austria, October 27–29, 2010. Proceedings, volume 6430 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16566-4 (print), 3-642-16567-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16567-2.

Bouchon-Meunier:1991:UKB

[BMYZ91] B. Bouchon-Meunier, Ronald R. Yager, and Lotfi Asker Zadeh, editors. Uncertainty in knowledge bases: 3rd International Conference on Information Processing and Management of Uncertainty in Knowledge-Based Systems, IPMU ’90, Paris, France, July 2–6, 1990, proceedings, volume 521 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54346-5 (Berlin), 0-387-54346-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q375 .I58 1990.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0521.
htm;   http://www.springerlink.
com/content/978-3-540-
54346-6;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=521.

Bouchon-Meunier:1995:AIC

[BMYZ95] B. Bouchon-Meunier, Ronald R. Yager, and Lotfi Asker Zadeh, editors. Advances in intelligent computing — IPMU ’94: 5th International Conference on Information Processing and Management of Uncertainty in Knowledge-Based Systems, Paris, France, July 1994: proceedings, volume 945 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60116-3 (Berlin: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.945.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0945.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60116-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=945.

Bieri:1991:CGM

[BN91a] H. Bieri and Hartmut Noltemeier, editors. Computational geometry: methods, algorithms, and applications: International Workshop on Computational Geometry CG ’91, Bern, Switzerland, March 21–22, 1991: proceedings, volume 553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54891-2 (Berlin), 0-387-54891-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA448.D38 I58 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0553.htm;
 http://www.springerlink.
com/content/978-3-540-
54891-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=553.

Biswas:1991:FST

[BN91b] S. Biswas and Kesav V.

Nori, editors. Foundations of software technology and theoretical computer science: 11th conference, New Delhi, India, December 17– 19, 1991: proceedings, volume 560 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54967-6 (Berlin), 0-387-54967-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F68 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0560.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54967-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=560.

Barendregt:1994:TPP

[BN94] H. P. Barendregt and Tobias Nipkow, editors. Types for proofs and programs: International Workshop TYPES ’93, Nijmegen, The Netherlands, May 24–28, 1993: Selected Papers, volume 806 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-58085-9 (New York: paperback), 3-540-58085-9 (Berlin: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T97 1994. DM72.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0806.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        58085-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=806.

Burkhardt:1998:CVEa

[BN98a] H. (Hans) Burkhardt and Bernd Neumann, editors. Computer vision — ECCV’98: 5th European Conference on Computer Vision, Freiburg, Germany, June 2–6, 1998: proceedings. Volume I, volume 1406 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64569-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1406-1407, TA1634 .E97 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1406.htm;
http://www.springerlink.
com/content/978-3-540-
64569-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1406.

Burkhardt:1998:CVEb

[BN98b] H. (Hans) Burkhardt and Bernd Neumann, editors. Computer vision — ECCV’98: 5th European Conference on Computer Vision, Freiburg, Germany, June 2–6, 1998: proceedings. Volume II, volume 1407 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64613-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634.E97 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1407.
htm;   http://www.springerlink.
com/content/978-3-540-
64613-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1407.

Bienstock:2004:IPC

[BN04] Daniel Bienstock and George Nemhauser, editors. Integer Programming and Combinatorial Optimization: 10th International IPCO Conference, New York, NY, USA, June 7–11, 2004: Proceedings, volume 3064 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22113-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T57.74 .C625 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3064.htm;
 http://www.springerlink.
com/content/978-3-540-
22113-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

0302-9743&volume=3064;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97946.

Backes:2009:CSE

[BN09a] Michael Backes and Peng Ning, editors. Computer Security — ES-ORICS 2009: 14th European Symposium on Research in Computer Security, Saint-Malo, France, September 21– 23, 2009. Proceedings, volume 5789 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04443-3 (print), 3-642-04444-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04444-1.

Boyd:2009:ISP

[BN09b] Colin Boyd and Juan González Nieto, editors. Information Security and Privacy: 14th Australasian Conference, ACISP 2009 Brisbane, Australia, July 1–3, 2009 Proceedings, volume 5594 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02619-2 (print), 3-642-02620-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02620-1.

Beynon:2001:CTI

[BND01] Meurig Beynon, Chrystopher L. Nehaniv, and Kerstin Dautenhahn, editors. Cognitive technology: instruments of mind: 4th International Conference, CT 2001, Coventry, UK, August 6–9, 2001: Proceedings, volume 2117 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42406-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I57 2001; QA267.A1 L43 no.2117. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2117.htm;
 http://www.springerlink.
com/content/978-3-540-
42406-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2117.

Borgefors:2000:DGC

[BNS00] Gunilla Borgefors, Ingela Nyström, and Gabriella Sanniti di Baja, editors. Discrete geometry for computer imagery: 9th International Conference, DGCI 2000, Uppsala, Sweden, December 13–15, 2000: Proceedings, volume 1953 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41396-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .I516 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1953.
htm;   http://www.springerlink.
com/content/978-3-540-
41396-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1953.

Berghofer:2009:TPH

[BNUW09] Stefan Berghofer, Tobias Nipkow, Christian Urban, and Makarius Wenzel, editors. Theorem Proving in Higher Order Logics: 22nd International Conference, TPHOLs 2009, Munich, Germany, August 17– 20, 2009. Proceedings, volume 5674 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03358-X (print), 3-642-03359-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03359-9.

Bravetti:2006:WSF

[BNZ06] Mario Bravetti, Manuel Núñez, and Gianluigi Zavattaro, editors. Web Services and Formal Methods: Third International Workshop, WS-FM 2006 Vienna, Austria, September 8– 9, 2006 Proceedings, volume 4184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38862-1 (print), 3-540-38865-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        38865-4.

Bjorner:1980:TFD

[BO80] D. Bjorner and O. N. Oest. Towards a formal description of Ada, volume 98 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10283-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 630 pp. LCCN QA76.73.A35 T68 1980. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0098.htm;
 http://www.springerlink.
com/content/978-0-387-
10283-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=98.

“References and bibliography”: p. 626–630.

Backhouse:2000:MPC

[BO00a] Roland C. Backhouse and Jose N. Oliveira, editors. Mathematics of program construction: 5th International Conference, MPC 200, Ponte de Lima, Portugal, July 200: Proceedings, volume 1837 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67727-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1837. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t1837.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67727-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1837.

Bomarius:2000:PFS

[BO00b] Frank Bomarius and Markku Oivo, editors. Product focused software process improvement: Second International Conference PRO-FES 2000, Oulu, Finland, June 20–22, 2000: Proceedings, volume 1840 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67688-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I5743 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1840.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67688-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1840.

Bockle:1995:EFG

[Boc95] Gunther Bockle. Exploitation of fine-grain parallelism, volume 942 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60054-X. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 188 pp. LCCN QA76.58 .B63 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0942.htm;
 http://www.springerlink.
com/content/978-3-540-
60054-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=942.

Bode:1991:DMC

[Bod91] Arndt Bode, editor. Distributed memory computing: 2nd European Conference, EDMCC2 Munich, FRG, April 22–24, 1991: proceedings, volume 487 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-53951-4, 0-387-53951-4 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .E96 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0487.htm;
 http://www.springerlink.
com/content/978-3-540-
53951-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=487.

Bode:1996:PVM

[Bod96] Arndt Bode, editor. Parallel virtual machine, EuroPVM ’96: third European PVM conference, Munich, Germany, October 7–9, 1996: proceedings, volume 1156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E975 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1156.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61779-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1156.

Bodlaender:2003:GTC

[Bod03] Hans L. Bodlaender, editor. Graph-Theoretic Concepts in Computer Science: 29th International Workshop, WG 2003, Elspeet, The Netherlands, June 2003: Revised Papers, volume 2880 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20452-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2880.htm;
 http://www.springerlink.
com/content/978-3-540-
20452-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2880;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b93953.

Bodik:2005:CCI

[Bod05] Rastislav Bodik, editor.

Compiler Construction: 14th International Conference, CC 2005, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2005, Edinburgh, UK, April 4–8, 2005: Proceedings, volume 3443 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25411-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 C32 2005.

URL http://www.springerlink.
com/content/978-3-540-
25411-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3443;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b107108.

Boers:2001:AEC

[Boe01] Egbert J. W. Boers, editor. Applications of evolutionary computing: EvoWorkshops 2001: EvoCOP, EvoFlight, EvoIASP, EvoLearn, and EvoSTIM, Como, Italy, April 18–20, 2001: Proceedings, volume 2037 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41920-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E899 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2037.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41920-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2037.

Bohm:1975:LCC

[Boh75] C. Bohm, editor. [Lambda]- calculus and computer science theory: proceedings of the symposium held in Rome, March 25– 27, 1975, volume 37 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CO-DEN LNCSD9. ISBN 0-387-07416-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.5.L351. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0037.htm;
 http://www.springerlink.
com/content/978-0-387-
07416-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=37.

English or French.

Bui:2012:SEL

[BOH^{+}12] Lam Thu Bui, Yew Soon Ong, Nguyen Xuan Hoai, Hisao Ishibuchi, and Ponnuthurai Nagaratnam Suganthan, editors. Simulated Evolution and Learning: 9th International Conference, SEAL 2012, Hanoi, Vietnam, December 16–19, 2012. Proceedings, volume 7673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34858-0 (print), 3-642-34859-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        34859-4.

Bolc:1978:NLC

[Bol78]  Leonard Bolc.  Nat-

ural language communication with computers, volume 63 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1978. CO-DEN LNCSD9. ISBN 0-387-08911-X. ISSN 0302-9743 (print), 1611-3349 (electronic). 292 pp. LCCN QA76.9 .I58N37. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0063.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        08911-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=63.

Boley:1999:TPI

[Bol99] H. (Harold) Boley. A tight, practical integration of relations and functions, volume 1712 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66644-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 169 pp. LCCN QA76.63 .B65 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1712.htm;
 http://www.springerlink.
com/content/978-3-540-
66644-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1712.

Boneh:2003:ACC

[Bon03] Dan Boneh, editor. Advances in Cryptology— CRYPTO 2003: 23rd Annual International Cryptology Conference Santa Barbara, California, USA, August 17– 21, 2003 Proceedings, volume 2729 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40674-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2729.
htm;   http://www.springerlink.
com/content/978-3-540-
40674-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2729;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11817.

Book:1991:RTA

[Boo91] Ronald V. Book, editor. Rewriting techniques and applications: 4th International Conference, RTA-91, Como, Italy, April 10–12, 1991: proceedings, volume 488 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-53904-2 (U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I555 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0488.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        53904-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=488.

Borger:1987:CTL

[Bor87] E. Borger. Computation theory and logic, volume 270 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-18170-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 442 pp. LCCN QA267 .C588 1987. URL http:

                                        //link.springer-ny.

com/link/service/series/
0558/tocs/t0270.htm;
 http://www.springerlink.
com/content/978-0-387-
18170-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=270.

Borger:1991:CSL

[Bor91] E. Borger, editor. Computer science logic: 4th workshop, CSL ’90, Heidelberg, Germany, October 1–5, 1990: proceedings, volume 533 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54487-9 (Berlin), 0-387-54487-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W624 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0533.htm;
 http://www.springerlink.
com/content/978-3-540-
54487-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=533.

A selection of papers presented at the workshop CSL’90 which was held at the Max-Planck-Haus in Heidelberg, Germany from October 1–5, 1990.

Borger:1992:CSL

[Bor92] E. Borger, editor. Computer science logic: 5th Workshop, CSL ’91, Berne, Switzerland, October 1991: proceedings, volume 626 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55789-X (Berlin), 0-387-55789-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W624 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0626.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55789-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=626.

Borger:1993:CSL

[Bor93] E. Borger, editor. Computer science logic: 6th workshop, CSL ’92, San Miniato, Italy, September 28–October 2, 1992: Selected Papers, volume 702 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56992-8 (Berlin), 0-387-56992-8 (New York).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W624 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0702.htm;
http://www.springerlink.
com/content/978-3-540-
56992-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=702.

Borning:1994:PPC

[Bor94] Alan Borning, editor.

Principles and practice of constraint programming: second International Workshop, PPCP ’94, Rosario, Orcas Island, WA, USA, May 2–4, 1994: proceedings, volume 874 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-58601-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .W67 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0874.htm;
http://www.springerlink.
com/content/978-0-387-
58601-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=874.

Boszormenyi:1996:PCT

[Bos96] Laszlo Boszormenyi, editor. Parallel computation: Third International ACPC Conference with special emphasis on parallel databases and parallel I/O, Klagenfurt, Austria, September 23–25, 1996: proceedings, volume 1127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61695-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I5 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1127.
htm;   http://www.springerlink.
com/content/978-3-540-
61695-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1127.

Bosma:2000:ANT

[Bos00a] Wieb Bosma, editor. Algorithmic number theory: 4th International Symposium, ANTS-IV, Leiden, the Netherlands, July 2–7, 2000: Proceedings, volume 1838 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67695-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1838. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1838.htm;
 http://www.springerlink.
com/content/978-3-540-
67695-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1838.

Bossi:2000:LBP

[Bos00b] Annalisa Bossi, editor.

Logic-based program synthesis and transformation: 9th International Workshop, LOPSTR’99, Venice, Italy, September 22–24, 1999: Selected Papers, volume 1817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67628-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1817. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1817.htm;
 http://www.springerlink.
com/content/978-3-540-
67628-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1817.

Bosch:2001:GCB

[Bos01] Jan Bosch, editor. Generative and componentbased software engineering: Third International Conference, GCSE 2001, Erfurt, Germany, September 10–13, 2001: Proceedings, volume 2186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42546-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .G418 2001; QA267.A1 L43 no.2186. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2186.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42546-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2186.

Bouchachia:2011:AIS

[Bou11] Abdelhamid Bouchachia, editor. Adaptive and Intelligent Systems: Second International Conference, ICAIS 2011, Klagenfurt, Austria, September 6–8, 2011. Proceedings, volume 6943 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23856-4 (print), 3-642-23857-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23857-4.

Bowers:1994:DDB

[Bow94] D. S. Bowers, editor. Directions in databases: 12th British National Conference on Databases, BNCOD 12, Guildford, United Kingdom, July 6–8, 1994: proceedings, volume 826 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58235-5 (Berlin), 0-387-58235-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B75 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0826.htm;
http://www.springerlink.
com/content/978-3-540-
58235-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=826.

Boyd:1995:CCI

[Boy95] Colin Boyd, editor.

Cryptography and coding: 5th IMA conference, Cirencester, UK, December 18–20, 1995, proceedings, volume 1025 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60693-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .C76 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1025.htm;
 http://www.springerlink.
com/content/978-3-540-
60693-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1025.

Papers from the 5th IMA Conference on Cryptography and Coding.

Boyd:2001:ACA

[Boy01] Colin Boyd, editor. Advances in cryptology — ASIACRYPT 2001: 7th International Conference on the Theory and Application of Cryptology and Information Security, Gold Coast, Australia, December 9–13, 2001: Proceedings, volume 2248 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42987-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I555 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2248.htm;
http://www.springerlink.
com/content/978-3-540-
42987-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2248.

Bruynooghe:1993:PLI

[BP93] M. Bruynooghe and Jaan Penjam, editors. Programming language implementation and logic programming: 5th International Symposium, PLILP ’93, Tallinn, Estonia, August 1993: proceedings, volume 714 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-57186-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .I5773 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0714.htm;
http://www.springerlink.
com/content/978-0-387-
57186-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=714.

Bosselaers:1995:IPS

[BP95] Antoon Bosselaers and Bart Preneel. Integrity primitives for secure information systems: final RIPE report of RACE Integrity Primitives Evaluation (R1040), volume 1007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60640-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). 239 pp. LCCN QA76.9.A25 I553 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1007.htm;
 http://www.springerlink.
com/content/978-3-540-
60640-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1007.

Blanzieri:2001:ACB

[BP01] Enrico Blanzieri and Luigi Portinale, editors. Advances in case-based reasoning: 5th European workshop, EWCBR 2000, Trento, Italy, September 6–9, 2000: Proceedings, volume 1898 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-67933-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1898; QA76.76.E95 E935 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1898.
htm;   http://www.springerlink.
com/content/978-3-540-
67933-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1898.

Benson:2003:ABT

[BP03] Gary Benson and Roderic Page, editors. Algorithms in Bioinformatics: Third International Workshop, WABI 2003, Budapest, Hungary, September 2003: Proceedings, volume 2812 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20076-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .W32 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2812.htm;
 http://www.springerlink.
com/content/978-3-540-
20076-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2812;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b13243.

Barakat:2004:PAN

[BP04a] Chadi Barakat and Ian Pratt, editors. Passive and Active Network Measurement: 5th International Workshop, PAM 2004, Antibes Juan-les-Pins, France, April 2004: Proceedings, volume 3015 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21492-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3015.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21492-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3015;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96961.

Bouge:2004:HPC

[BP04b] L. (Luc) Bougé and V. K. Prasanna Kumar, editors. High performance computing — HiPC 2004: 11th international conference, Bangalore, India, December 19–22, 2004: proceedings, volume 3296 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24129-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I566 2004. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3296;   http:
//www.springerlink.
com/content/978-3-540-
24129-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3296.

Borrione:2005:CHD

[BP05] Dominique Borrione and Wolfgang Paul, editors. Correct hardware design and verification methods: 13th IFIP WG 10.5 advanced research working conference, Charme 2005, Saarbruecken, Germany, October 3–6, 2005, proceedings, volume 3725 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29105-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        29105-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3725.

Basili:2007:AAI

[BP07] Roberto Basili and Maria Teresa Pazienza, editors. AI*IA 2007: Artificial Intelligence and Human-Oriented Computing: 10th Congress of the Italian Association for Artificial Intelligence, Rome, Italy, September 10– 13, 2007. Proceedings, volume 4733 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74781-8 (print), 3-540-74782-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74782-6.

Bensalem:2009:RVI

[BP09a] Saddek Bensalem and Doron A. Peled, editors. Runtime Verification: 9th International Workshop, RV 2009, Grenoble, France, June 26–28, 2009. Selected Papers, volume 5779 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04693-2 (print), 3-642-04694-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04694-0.

Bournez:2009:RPI

[BP09b] Olivier Bournez and Igor Potapov, editors. Reachability Problems: 3rd International Workshop, RP 2009, Palaiseau, France, September 23– 25, 2009. Proceedings, volume 5797 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04419-0 (print), 3-642-04420-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04420-5.

Bellatreche:2011:MDE

[BP11] Ladjel Bellatreche and Filipe Mota Pinto, editors. Model and Data Engineering: First International Conference, MEDI 2011, Óbidos, Portugal, September 28–30, 2011. Proceedings, volume 6918 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24442-4 (print), 3-642-24443-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24443-8.

Barcelo:2012:DAI

[BP12a] Pablo Barceló and Reinhard Pichler, editors. Datalog in Academia and Industry: Second International Workshop, Datalog 2.0, Vienna, Austria, September 11– 13, 2012. Proceedings, volume 7494 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32924-1 (print), 3-642-32925-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32925-8.

Brorsson:2012:RST

[BP12b] Mats Brorsson and Luís Miguel Pinho, editors. Reliable Software Technologies — Ada-Europe 2012: 17th Ada-Europe International Conference on Reliable Software Technologies, Stockholm, Sweden, June 11– 15, 2012. Proceedings, volume 7308 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30597-0 (print), 3-642-30598-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30598-6.

Baranauskas:2007:HCIa

[BPAB07a] Cécilia Baranauskas, Philippe Palanque, Julio Abascal, and Simone Diniz Junqueira Barbosa, editors. Human-Computer Interaction — INTERACT 2007: 11th IFIP TC 13 International Conference, Rio de Janeiro, Brazil, September 10– 14, 2007, Proceedings, Part I, volume 4662 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74794-X (print), 3-540-74796-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        74796-3.

Baranauskas:2007:HCIb

[BPAB07b] Cécilia Baranauskas, Philippe Palanque, Julio Abascal, and Simone Diniz Junqueira Barbosa, editors. Human-Computer Interaction — INTERACT 2007: 11th IFIP TC 13 International Conference, Rio de Janeiro, Brazil, September 10– 14, 2007, Proceedings, Part II, volume 4663 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74799-0 (print), 3-540-74800-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74800-7.

Barker-Plummer:2006:DRI

[BPCS06] Dave Barker-Plummer, Richard Cox, and Nik Swoboda, editors. Diagrammatic Representation and Inference: 4th International Conference, Diagrams 2006, Stanford, CA, USA, June 28–30, 2006. Proceedings, volume 4045 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35623-1 (print), 3-540-35624-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35624-0.

Boissier:2006:COI

[BPD06] Olivier Boissier, Julian Padget, and Virginia Dignum, editors. Coordination, Organizations, Institutions, and Norms in Multi-Agent Systems: AAMAS 2005 International Workshops on Agents, Norms and Institutions for Regulated Multiagent Systems, ANIREM 2005, and Organizations in Multi-Agent Systems, OOOP 2005, Utrecht, The Netherlands, July 25–26, 2005, Revised Selected Papers, volume 3913 of Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35173-6 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35173-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3913.

Barahona:1991:EPC

[BPP91] P. Barahona, Luis Moniz Pereira, and A. Porto, editors. EPIA ’91: 5th Portuguese Conference on Artificial Intelligence, Albufeira, Portugal, October 1–3, 1991: proceedings, volume 541 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54535-2 (Berlin), 0-387-54535-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P67 1991.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0541.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        54535-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=541.

Bellahsene:2002:OOI

[BPR02] Zohra Bellahsène, Dilipkumar Patel, and Colette Rolland, editors. Object-Oriented. Information Systems: 8th International Conference, OOIS 2002, Montpellier, France, September 2–5, 2002. Proceedings, volume 2425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44087-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35 I52 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2425.htm;
 http://www.springerlink.
com/content/978-3-540-
44087-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2425.

Bastide:2005:EHC

[BPR05] Rémi Bastide, Philippe Palanque, and Jörg Roth, editors. Engineering Human Computer Interaction and Interactive Systems: Joint Working Conferences EHCI-DSVIS 2004, Hamburg, Germany, July 11–13, 2004, Revised Selected Papers, volume 3425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26097-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 E44 2005.

URL http://www.springerlink.
com/content/978-3-540-
26097-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3425.

Becker:2008:QSA

[BPR08] Steffen Becker, Frantisek Plasil, and Ralf Reussner, editors. Quality of Software Architectures. Models and Architectures: 4th International Conference on the Quality of Software-Architectures, QoSA 2008, Karlsruhe, Germany, October 14–17, 2008. Proceedings, volume 5281 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87878-5 (print), 3-540-87879-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87879-7.

Butler:2002:IFM

[BPS02] Michael Butler, Luigia Petre, and Kaisa Sere, editors. Integrated formal methods: Third International Conference, IFM 2002, Turku, Finland, May 15–18, 2002: Proceedings, volume 2335 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43703-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 I58 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2335.htm;
http://www.springerlink.
com/content/978-3-540-
43703-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2335.

Boavida:2006:NNT

[BPS06] Fernando Boavida, Thomas Plagemann, and Burkhard Stiller, editors. Networking 2006. Networking Technologies, Services, and Protocols; Performance of Computer and Communication Networks; Mobile and Wireless Communications Systems: 5th International IFIP-TC 6 Networking Conference, Coimbra, Portugal, May 15–19, 2006, Proceedings, volume 3976 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34192-7 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34192-5.

Barthe:2011:SEF

[BPS11] Gilles Barthe, Alberto Pardo, and Gerardo Schneider, editors. Software Engineering and Formal Methods: 9th International Conference, SEFM 2011, Montevideo, Uruguay, November 14–18, 2011. Proceedings, volume 7041 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24689-3 (print), 3-642-24690-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24690-6.

Bader:2006:HPC

[BPSP06] David A. Bader, Manish Parashar, Varadarajan Sridhar, and Viktor K. Prasanna, editors. High performance computing — HIPC 2005: 12th international conference, Goa, India, December 18–21, 2005. proceedings, volume 3769 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30936-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
30936-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3769.

Bugliesi:2006:ALPa

[BPSW06a] Michele Bugliesi, Bart Preneel, Vladimiro Sassone, and Ingo Wegener, editors. Automata, Languages and Programming: 33rd International Colloquium, ICALP 2006, Venice, Italy, July 10–14, 2006, Proceedings, Part I, volume 4051 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35904-4 (print), 3-540-35905-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35905-0.

Bugliesi:2006:ALPb

[BPSW06b] Michele Bugliesi, Bart Preneel, Vladimiro Sassone, and Ingo Wegener, editors. Automata, Languages and Programming: 33rd International Colloquium, ICALP 2006, Venice, Italy, July 10–14, 2006, Proceedings, Part II, volume 4052 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35907-9 (print), 3-540-35908-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        35908-1.

Bloch:2006:FLA

[BPT06] Isabelle Bloch, Alfredo Petrosino, and Andrea G. B. Tettamanzi, editors. Fuzzy Logic and Applications: 6th International Workshop, WILF 2005, Crema, Italy, September 15–17, 2005, Revised Selected Papers, volume 3849 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32529-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32529-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3849.

Becker:2004:FPL

[BPV04] Jürgen Becker, Marco Platzner, and Serge Vernalde, editors. Field Programmable Logic and Application: 14th International Conference, FPL 2004, Leuven, Belgium, August 30–September 1, 2004, Proceedings, volume 3203 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22989-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3203.htm;
 http://www.springerlink.
com/content/978-3-540-
22989-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3203;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99787.

Bauknecht:2006:CWT

[BPW06] Kurt Bauknecht, Birgit Pröll, and Hannes Werthner, editors. E-Commerce and Web Technologies: 7th International Conference, EC-Web 2006, Krakow, Poland, September 5– 7, 2006. Proceedings, volume 4082 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37743-3 (print), 3-540-37745-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37745-0.

Bernardo:2009:FMW

[BPZ09] Marco Bernardo, Luca Padovani, and Gianluigi Zavattaro, editors. Formal Methods for Web Services: 9th International School on Formal Methods for the Design of Computer, Communication, and Software Systems, SFM 2009, Bertinoro, Italy, June 1– 6, 2009, Advanced Lectures, volume 5569 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01917-X (print), 3-642-01918-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01918-0.

Bermond:1989:DAI

[BR89] J.-C. Bermond and M. Raynal, editors. Distributed algorithms: 3rd International Workshop, Nice, France, September 26– 28, 1989: proceedings, volume 392 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51687-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D4852 1989. DM47.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0392.htm;
 http://www.springerlink.
com/content/978-0-387-
51687-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=392.

Proceedings from the Third International Workshop on Distributed Algorithms held at La Colle-sur-Loup, Sept. 26–28, 1989.

Belli:1992:IEA

[BR92] Fevzi Belli and Franz-Josef Radermacher, editors. Industrial and engineering applications of artificial intelligence and expert systems: 5th International Conference, IEA/AIE-92, Paderborn, Germany, June 9–12, 1992: proceedings, volume 604 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55601-X (Berlin), 0-387-55601-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.604. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0604.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55601-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=604.

Proceedings of the Fifth International Conference on Industrial and Engineering Applications of Artificial Intelligence and Expert Systems (IEA/AIE-92).

Billington:1996:ATP

[BR96a] Jonathan Billington and Wolfgang Reisig, editors. Application and theory of Petri Nets, 1996: 17th International Conference, Osaka, Japan, June 24–28, 1996: proceedings, volume 1091 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61363-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I56 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1091.htm;
 http://www.springerlink.
com/content/978-3-540-
61363-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1091.

Burke:1996:PTA

[BR96b] Edmund Burke and Peter (Peter Malcolm) Ross, editors. Practice and theory of automated timetabling: first international conference, Edinburgh, UK, August 29–September 1, 1995: Selected Papers, volume 1153 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61794-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q341 .P7 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1153.htm;
 http://www.springerlink.
com/content/978-3-540-
61794-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1153.

Bertran:1997:TBR

[BR97] Miquel Bertran and Teodor Rus, editors. Transformation-based reactive systems development: 4th International AMAST Workshop on Real-Time Systems and Concurrent and Distributed Software, ARTS’97, Palma, Mallorca, Spain, May 21–23, 1997: proceedings, volume 1231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63010-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54 .I54 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1231.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63010-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1231.

Broy:1998:RTS

[BR98] Manfred Broy and Bernhard Rumpe, editors. Requirements targeting software and systems engineering: International Workshop RTSE ’97, Bernried, Germany, October 12–14, 1997: Proceedings, volume 1526 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65309-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758.I578 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1526.
htm;   http://www.springerlink.
com/content/978-3-540-
65309-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1526.

Bittencourt:2002:AAI

[BR02] Guilherme Bittencourt and Geber L. Ramalho, editors. Advances in Artificial Intelligence: 16th Brazilian Symposium on Artificial Intelligence, SBIA 2002 Porto de Galinhas/Recife, Brazil, November 11–14, 2002. Proceedings, volume 2507 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00124-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .S5434 2002; Q334 .S56 2002. URL

http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t2507.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        00124-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2507.

Also available via the

World Wide Web.

Ball:2003:MCS

[BR03] Thomas Ball and Sriram K. Rajamani, editors. Model checking software: 10th International SPIN Workshop, Portland, OR, USA, May 9–10, 2003: Proceedings, volume 2648 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40117-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 S65 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2648.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40117-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2648.

Also available via the

World Wide Web.

Basin:2004:ARS

[BR04] David Basin and Michael Rusinowitch, editors. Automated Reasoning: Second International Joint Conference, IJCAR 2004, Cork, Ireland, July 4– 8, 2004. Proceedings, volume 3097 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22345-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3097.htm;
 http://www.springerlink.
com/content/978-3-540-
22345-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3097;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98691.

Bengio:2006:MLM

[BR06] Samy Bengio and Steve Renals, editors. Machine learning for multimodal interaction: second international workshop, MLMI 2005, Edinburgh, UK, July 11– 13, 2005: revised selected papers, volume 3869 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32549-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .W69 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32549-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3869.

Botana:2007:ADG

[BR07a] Francisco Botana and Tomas Recio, editors. Automated Deduction in Geometry: 6th International Workshop, ADG 2006, Pontevedra, Spain, August 31–September 2, 2006. Revised Papers, volume 4869 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77355-X (print), 3-540-77356-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77356-6.

Bozapalidis:2007:AIS

[BR07b] Symeon Bozapalidis and George Rahonis, editors. Algebraic Informatics: Second International Conference, CAI 2007, Thessaloniki, Greece, May 21–25, 2007, Revised Selected and Invited Papers, volume 4728 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75413-X (print), 3-540-75414-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75414-5.

Burke:2007:PTA

[BR07c] Edmund K. Burke and Hana Rudová, editors. Practice and Theory of Automated Timetabling VI: 6th International Conference, PATAT 2006 Brno, Czech Republic, August 30–September 1, 2006 Revised Selected Papers, volume 3867 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77344-4 (print), 3-540-77345-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77345-0.

Bozapalidis:2009:AIT

[BR09] Symeon Bozapalidis and George Rahonis, editors.

Algebraic Informatics: Third International Conference, CAI 2009, Thessaloniki, Greece, May 19–22, 2009, Proceedings, volume 5725 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03563-9 (print), 3-642-03564-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03564-7.

Bang:2012:PTD

[BR12a] Magnus Bang and Eva L. Ragnemalm, editors. Persuasive Technology. Design for Health and Safety: 7th International Conference, PERSUA-SIVE 2012, Linköping, Sweden, June 6–8, 2012. Proceedings, volume 7284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31036-2 (print), 3-642-31037-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31037-9.

Bodenreider:2012:DIL

[BR12b] Olivier Bodenreider and Bastien Rance, editors. Data Integration in the Life Sciences: 8th International Conference, DILS 2012, College Park, MD, USA, June 28–29, 2012. Proceedings, volume 7348 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31039-7 (print), 3-642-31040-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31040-9.

Brauer:1973:DJH

[Bra73] Wilfried Brauer, editor. 3. [i.e., dritte] Jahrestagung: Hamburg, 8.–10. Oktober 1973 [Third Annual Conference: Hamburg, 8–10 October 1973], volume 1 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1973. CO-DEN LNCSD9. ISBN 3-540-06473-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t0001.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        06473-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1. En-

glish or German.

Brakhage:1975:ATF

[Bra75] H. Brakhage, editor. Automata theory and formal languages: 2nd GI conference, Kaiserslautern, May 20–23, 1975, volume 33 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07407-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A9241. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0033.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        07407-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=33.

English or German.

Branquart:1976:OTP

[Bra76] Paul Branquart. An Optimized translation process and its application to ALGOL 68, volume 38 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1976. CO-DEN LNCSD9. ISBN 0-387-07545-3. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 334 pp. LCCN QA76.73.A24 O67. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0038.htm;
 http://www.springerlink.
com/content/978-0-387-
07545-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=38.

Brauer:1980:NTA

[Bra80] Wilfried Brauer, editor.

Net theory and applications: proceedings of the Advance Course on General Net Theory of Processes and Systems, Hamburg, October 8– 19, 1979, volume 84 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10001-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .N47 1980. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0084.htm;
 http://www.springerlink.
com/content/978-0-387-
10001-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=84.

Brauer:1985:ALP

[Bra85] Wilfried Brauer, editor. Automata, languages, and programming: 12th colloquium, Nafplion, Greece, July 15–19, 1985, volume 194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15650-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.194. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0194.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        15650-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=194.

Proceedings of the 12th International Colloquium on Automata, Languages, and Programming, sponsored by the European Association for Theoretical Computer Science and organized by the National Technical University of Athens with the University of Hamburg.

Brassard:1988:MCT

[Bra88] Gilles Brassard. Modern cryptology: a tutorial, volume 325 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-96842-3. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 107 pp. LCCN Z103 .B721 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0325.htm;
 http://www.springerlink.
com/content/978-0-387-
96842-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=325.

Brassard:1990:ACC

[Bra90] Gilles Brassard, editor.

Advances in cryptology: CRYPTO ’89: proceedings, volume 435 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-97317-6, 3-540-97317-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0435.htm;
 http://www.springerlink.
com/content/978-0-387-
97317-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=435.

Conference held Aug. 20– 24, 1989 at the University of California, Santa Barbara.

Brazdil:1993:MLE

[Bra93] Pavel B. Brazdil, editor. Machine learning, ECML-93: European Conference on Machine Learning, Vienna, Australia, April 5–7, 1993: proceedings, volume 667 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56602-3 (New York), 3-540-56602-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 1993. DM86.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0667.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        56602-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=667.

Brandenburg:1996:GDS

[Bra96] F. J. (Franz J.) Brandenburg, editor. Graph drawing: Symposium on Graph Drawing, GD ’95, Passau, Germany, September 20–22, 1995: proceedings, volume 1027 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60723-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1027.
htm;   http://www.springerlink.
com/content/978-3-540-
60723-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1027.

Braffort:1999:GBC

[Bra99] Anneliese Braffort, editor. Gesture-based communication in human computer interaction: International Gesture Workshop, GW ’99, Gif-sur-Yvette, France, March 17–19, 1999: Proceedings, volume 1739 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66935-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I618 1999.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t1739.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66935-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1739.

Bradfield:2002:CSL

[Bra02] J. C. Bradfield, editor. Computer Science Logic: 16th International Workshop, CSL 2002, 11th Annual Conference of the EACSL, Edinburgh, Scotland, UK, September 22–25, 2002. Proceedings, volume 2471 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44240-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W624 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2471.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44240-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2471.

Branco:2007:AAA

[Bra07] António Branco, editor. Anaphora: Analysis, Algorithms and Applications: 6th Discourse Anaphora and Anaphor Resolution Colloquium, DAARC 2007, Lagos, Portugal, March 29–30, 2007. Selected Papers, volume 4410 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71411-1 (print), 3-540-71412-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71412-5.

Bradley:2009:CPE

[Bra09a] Jeremy T. Bradley, editor. Computer Performance Engineering: 6th European Performance Engineering Workshop, EPEW 2009 London, UK, July 9–10, 2009 Proceedings, volume 5652 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02923-X (print), 3-642-02924-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02924-0.

Bramer:2009:AII

[Bra09b] Max Bramer, editor. Artificial Intelligence An International Perspective: An International Perspective, volume 5640 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03225-7 (print), 3-642-03226-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03226-4.

Breu:1991:AST

[Bre91] R. (Ruth) Breu. Algebraic specification techniques in object oriented programming environments, volume 562 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-54972-2, 3-540-54972-2 (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 228 pp. LCCN QA76.64 .B74 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0562.htm;
 http://www.springerlink.
com/content/978-0-387-
54972-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=562.

Originally presented as the author’s thesis (Ph. D.).

Brezany:1997:IOI

[Bre97] Peter Brezany. Input/ output intensive massively parallel computing: language support, automatic parallelization, advanced optimization, and runtime systems, volume 1220 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62840-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 288 pp. LCCN QA76.58.B749 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1220.htm;
http://www.springerlink.
com/content/978-3-540-
62840-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1220.

Bressan:2003:EEX

[Bre03] Stephane Bressan, editor. Efficiency and effectiveness of XML tools and techniques and data integration over the Web: VLDB 2002 workshop EEXTT and CAiSE 2002 workshop DIWeb: Revised Papers, volume 2590 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00736-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.H94 V58 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2590.
htm;   http://www.springerlink.
com/content/978-3-540-
00736-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2590.

Also available via the World Wide Web.

Brickell:1993:ACC

[Bri93] Ernest F. Brickell, editor. Advances in Cryptology — CRYPTO ’92: 12th Annual International Cryptology Conference, Santa Barbara, California, USA, August 16–20, 1992: Proceedings, volume 740 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57340-2 (New York), 3-540-57340-2 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1992. DM104.00. URL

http://link.springer-

ny.com/link/service/
series/0558/tocs/t0740.
htm;   http://www.springerlink.
com/content/978-0-387-
57340-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=740.

Brinksma:1995:TAC

[Bri95] Ed Brinksma, editor.

Tools and algorithms for the construction and analysis of systems: first International Workshop, TACAS ’95, Aarhus, Denmark, May 19–20, 1995: Selected Papers, volume 1019 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60630-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T66 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1019.htm;
http://www.springerlink.
com/content/978-3-540-
60630-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1019.

Brinksma:1997:TAC

[Bri97] Ed Brinksma, editor.

Tools and algorithms for the construction and analysis of systems: Third International Workshop, TACAS 97, Enschede, The Netherlands, April 2–4, 1997: proceedings, volume 1217 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62790-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T33 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1217.htm;
 http://www.springerlink.
com/content/978-3-540-
62790-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1217.

Bertolotto:2008:WWG

[BRL08] Michela Bertolotto, Cyril Ray, and Xiang Li, editors. Web and Wireless Geographical Information Systems: 8th International Symposium, W2GIS 2008, Shanghai, China, December 11–12, 2008. Proceedings, volume 5373 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89902-2 (print), 3-540-89903-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89903-7.

Boulicaut:2006:CBM

[BRM06] Jean-François Boulicaut, Luc Raedt, and Heikki Mannila, editors. Constraint-Based Mining and Inductive Databases: European Workshop on Inductive Databases and Constraint Based Mining, Hinterzarten, Germany, March 11–13, 2004, Revised Selected Papers, volume 3848 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31331-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31331-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3848.

Brookes:1992:MFP

[Bro92] S. D. Brookes, editor.

Mathematical foundations of programming semantics: 7th international conference, Pittsburgh, PA, USA, March 25–28, 1991: proceedings, volume 598 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55511-0 (Germany), 0-387-55511-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .M38 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0598.htm;
 http://www.springerlink.
com/content/978-3-540-
55511-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=598.

Proceedings of the Seventh International Conference on the Mathematical Foundations of Programming Semantics.

Brookes:1994:MFP

[Bro94] S. D. Brookes, editor. Mathematical foundations of programming semantics: 9th international conference, New Orleans, LA, USA, April 7–10, 1993: proceedings, volume 802 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58027-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.7 .M38 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0802.htm;
 http://www.springerlink.
com/content/978-0-387-
58027-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=802.

Proceedings of the Ninth International Conference on the Mathematical Foundations of Programming Semantics.

Brlek:2009:DGC

[BRP09] Srečko Brlek, Christophe Reutenauer, and Xavier Provençal, editors. Discrete Geometry for Computer Imagery: 15th IAPR International Conference, DGCI 2009, Montréal, Canada, September 30 — October 2, 2009. Proceedings, volume 5810 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04396-8 (print), 3-642-04397-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04397-0.

Brauer:1987:PNA

[BRR87a] Wilfried Brauer, Wolfgang Reisig, and Grzegorz Rozenberg, editors. Petri nets: applications and relationships to other models of concurrency: Advances in Petri nets 1986, part II: proceedings of an advanced course Bad Honnef, 8.-19. September 1986, volume 255 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-17906-2, 3-540-17906-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.255. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0255.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        17906-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=255.

Papers from the Advanced Course on Petri Nets held September 1986 in Bad Honnef, West Germany. Title on spine: Advances in Petri nets 1986, Part II.

Brauer:1987:PNC

[BRR87b] Wilfried Brauer, Wolfgang Reisig, and Grzegorz Rozenberg, editors. Petri nets, central models and their properties: advances in Petri nets 1986, part I: proceedings of an advanced course Bad Honnef, 8.-19. September 1986, volume 254 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17905-4 (paperback), 3-540-17905-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.254. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0254.htm;
http://www.springerlink.
com/content/978-0-387-
17905-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=254.

Papers from the Advanced Course on Petri Nets held September 1986 in Bad Honnef, West Germany. Title on spine: Advances in Petri nets 1986, Part I.

Bry:1997:DOO

[BRR97] François Bry, Raghu Ramakrishnan, and Kotagiri Ramamohanarao, editors. Deductive and object-oriented databases:

5th international conference, DOOD ’97, Montreux, Switzerland, December 1997: proceedings, volume 1341 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63792-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3I55826 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1341.htm;
 http://www.springerlink.
com/content/978-3-540-
63792-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1341.

Basili:2007:ESE

[BRS^{+}07] Victor R. Basili, Dieter Rombach, Kurt Schneider, Barbara Kitchenham, Dietmar Pfahl, and Richard W. Selby, editors. Empirical Software Engineering Issues. Critical Assessment and Future Directions: International Workshop, Dagstuhl Castle, Germany, June 26–30, 2006. Revised Papers, volume 4336 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71300-X (print), 3-540-71301-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71301-2.

Buth:2009:CSR

[BRS09] Bettina Buth, Gerd Rabe, and Till Seyfarth, editors. Computer Safety, Reliability, and Security: 28th International Conference, SAFECOMP 2009, Hamburg, Germany, September 15– 18, 2009. Proceedings, volume 5775 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04467-0 (print), 3-642-04468-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04468-7.

Blockeel:2008:ILP

[BRST08] Hendrik Blockeel, Jan Ramon, Jude Shavlik, and Prasad Tadepalli, editors. Inductive Logic Programming: 17th International Conference, ILP 2007, Corvallis, OR, USA, June 19–21, 2007, Revised Selected Papers, volume 4894 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78468-3 (print), 3-540-78469-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78469-2.

Brafman:2011:ADT

[BRT11] Ronen I. Brafman, Fred S. Roberts, and Alexis Tsoukiàs, editors. Algorithmic Decision Theory: Second International Conference, ADT 2011, Piscataway, NJ, USA, October 26–28, 2011. Proceedings, volume 6992 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24872-1 (print), 3-642-24873-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24873-3.

Bruynooghe:2004:LBP

[Bru04] Maurice Bruynooghe, editor. Logic Based Program Synthesis and Transformation: 13th International Symposium, LOPSTR 2003, Uppsala, Sweden, August 25–27, 2003: revised Selected Papers, volume 3018 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22174-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63.L69 2003. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3018.
htm;   http://www.springerlink.
com/content/978-3-540-
22174-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3018;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98080.

Boehme:2006:IIC

[BRU06a] Thomas Boehme, Victor M. Larios Rosillo, and Herwig Unger, editors. Innovative Internet Community Systems: 4th International Workshop, IICS 2004, Guadalajara, Mexico, June 21– 23, 2004. Revised Papers, volume 3473 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-28880-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28880-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3473.

Bruel:2006:SEM

[Bru06b] Jean-Michel Bruel, editor. Satellite Events at the MoDELS 2005 Conference: MoDELS 2005 International Workshops OCLWS, MoDeVA, MARTES, AOM, MTiP, WiSME, MODAUI, NfC, MDD, WUsCAM, Montego Bay, Jamaica, October 2–7, 2005, Revised Selected Papers, volume 3844 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31780-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        31780-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3844.

Brookes:1985:SCC

[BRW85] S. D. Brookes, A. W.

Roscoe, and G. Winskel, editors. Seminar on Concurrency: Carnegie-Mellon University, Pittsburgh, PA, July 9–11, 1984, volume 197 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15670-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S44941 1984. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0197.htm;
http://www.springerlink.
com/content/978-0-387-
15670-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=197.

Bode:1993:PPA

[BRW93] Arndt Bode, Mike Reeve, and Gottfried Wolf, editors. Parle ’93, parallel architectures and languages Europe: 5th International PARLE Conference, Munich, Germany, June 14–17, 1993: proceedings, volume 694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56891-3 (Berlin), 0-387-56891-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I564 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0694.htm;
 http://www.springerlink.
com/content/978-3-540-
56891-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=694.

Bauer:1976:LHI

[BS76] Friedrich Ludwig Bauer and Klaus Samelson, editors. Language hierarchies and interfaces: international summer school [held July 23 to August 2, 1975 in Marktoberdorf], volume 46 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1976. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L335.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0046.
htm;   http://www.springerlink.
com/content/978-0-387-
07994-3;   http://www.
springerlink.com/openurl.
asp?genre=book&isbn=
978-3-540-07994-1;

http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=46.

Berghammer:1992:GTC

[BS92] R. Berghammer and Gunther Schmidt, editors. Graph-theoretic concepts in computer science: 17th International Workshop WG ’91, Fischbachau, Germany, June 17–19, 1991, proceedings, volume 570 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55121-2 (Berlin), 0-387-55121-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I65 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0570.
htm;   http://www.springerlink.
com/content/978-3-540-
55121-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=570.

Borzyszkowski:1993:MFC

[BS93] Andrzej M. Borzyszkowski and S. Sokolowski, editors. Mathematical foundations of computer science 1993: 18th International Symposium, MFCS’93, Gdansk, Poland, August 30– September 3, 1993: proceedings, volume 711 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57182-5 (New York: softcover), 3-540-57182-5 (Berlin: softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 I54 1993. DM142.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0711.htm;
 http://www.springerlink.
com/content/978-0-387-
57182-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=711.

Bolding:1994:PCR

[BS94] Kevin Bolding and Lawrence Snyder, editors. Parallel computer routing and communication: first International Workshop, PCRCW ’94, Seattle, Washington, USA, May 16–18, 1994: proceedings, volume 853 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58429-3 (Berlin), 0-387-58429-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P39 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0853.htm;
http://www.springerlink.
com/content/978-3-540-
58429-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=853.

Bigun:2001:AVB

[BS01a] Josef Bigun and Fabrizio Smeraldi, editors. Audio-and video-based biometric person authentication: Third International Conference, AVBPA 2001, Halmstad, Sweden, June 6–8, 2001: Proceedings, volume 2091 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42216-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.S65 .A944 2001; QA267.A1 L43 no.2091.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2091.
htm;   http://www.springerlink.
com/content/978-3-540-
42216-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2091.

Boztas:2001:AAA

[BS01b] Serdar Boztaş and Igor E.

Shparlinski, editors. Applied Algebra, Algebraic Algorithms and Error-Correcting Codes: 14th International Symposium, AAECC-14, Melbourne, Australia, November 26–30, 2001: Proceedings, volume 2227 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42911-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2227.htm;
 http://www.springerlink.
com/content/978-3-540-
42911-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2227.

Buchsbaum:2001:AEE

[BS01c] Adam L. Buchsbaum and Jack Snoeyink, editors. Algorithm engineering and experimentation: Third International Workshop, ALENEX 2001, Washington, DC, USA, January 5–6, 2001: Revised Papers, volume 2153 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42560-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A42 2001; QA267.A1 L43 no.2153. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2153.htm;
 http://www.springerlink.
com/content/978-3-540-
42560-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2153.

Batten:2002:ISP

[BS02a] Lynn Margaret Batten and Jennifer Seberry, editors. Information Security and Privacy: 7th Australasian Conference, ACISP 2002, Melbourne, Australia, July 3–5, 2002: Proceedings, volume 2384 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43861-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 A278 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2384.
htm;   http://www.springerlink.
com/content/978-3-540-
43861-8;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2384.

Blieberger:2002:RST

[BS02b] Johann Blieberger and Alfred Strohmeier, editors. Reliable software technologies — Ada-Europe 2002: 7th Ada-Europe International Conference on reliable Software Technologies, Vienna, Austria, June 2002: Proceedings, volume 2361 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43784-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2361.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43784-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2361.

Benatallah:2003:TSI

[BS03a] Boualem Benatallah and Ming-Chien Shan, editors. Technologies for E-Services: 4th International Workshop, TES 2003, Berlin, Germany, September 7–8, 2003:

Proceedings, volume 2819 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20052-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88813 .T47 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2819.
htm;   http://www.springerlink.
com/content/978-3-540-
20052-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2819;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13474.

Boszormenyi:2003:MPL

[BS03b] László Böszörményi and Peter Schojer, editors. Modular Programming Languages: Joint Modular Languages Conference, JMLC 2003, Klagenfurt, Austria, August 25–27, 2003: proceedings: 2003, volume 2789 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40796-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 2003. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2789.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40796-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2789;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12023.

Bouyssounouse:2005:ESD

[BS05] Bruno Bouyssounouse and Joseph Sifakis, editors. Embedded systems design: the ARTIST roadmap for research and development, volume 3436 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25107-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 E64 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25107-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3436;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106761.

Beck:2006:IAT

[BS06a] J. Christopher Beck and Barbara M. Smith, editors. Integration of AI and OR Techniques in Constraint Programming for Combinatorial Optimization Problems: Third International Conference, CPAIOR 2006, Cork, Ireland, May 31– June 2, 2006, Proceedings, volume 3990 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34306-7 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34306-6.

Beichel:2006:CVA

[BS06b] Reinhard R. Beichel and Milan Sonka, editors. Computer Vision Approaches to Medical Image Analysis: Second International ECCV Workshop, CVAMIA 2006 Graz, Austria, May 12, 2006 Revised Papers, volume 4241 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46257-0 (print), 3-540-46258-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        46258-3.

Bunke:2006:DAS

[BS06c] Horst Bunke and A. Lawrence Spitz, editors. Document analysis systems VII: 7th international workshop, DAS 2006, Nelson, New Zealand, February 13–15, 2006: proceedings, volume 3872 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32140-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 I55 2006; TA1650 .D37 2006. URL

                                        http://www.springerlink.
                                        com/content/978-3-540-
                                        32140-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3872.

Bussler:2006:TSI

[BS06d] Christoph Bussler and Ming-Chien Shan, editors. Technologies for e-services: 6th international workshop, TES 2005, Trondheim, Norway, September 2–3, 2005: revised selected papers, volume 3811 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31067-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .T43 2005.

URL http://www.springerlink.
com/content/978-3-540-
31067-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3811.

Bampis:2009:AOA

[BS09] Evripidis Bampis and Martin Skutella, editors. Approximation and Online Algorithms: 6th International Workshop, WAOA 2008, Karlsruhe, Germany, September 18– 19, 2008. Revised Papers, volume 5426 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-93979-2 (print), 3-540-93980-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
93980-1.

Blaschke:2010:LLI

[BS10] Christian Blaschke and Hagit Shatkay, editors. Linking Literature, Information, and Knowledge for Biology: Workshop of the BioLink Special Interest Group, ISMB/ ECCB 2009, Stockholm, June 28–29, 2009, Revised Selected Papers, volume 6004 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13130-1 (print), 3-642-13131-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13131-8.

Bafna:2011:RCM

[BS11a] Vineet Bafna and S. Cenk Sahinalp, editors. Research in Computational Molecular Biology: 15th Annual International Conference, RECOMB 2011, Vancouver, BC, Canada, March 28–31, 2011. Proceedings, volume 6577 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20035-4 (print), 3-642-20036-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20036-6.

Balduccini:2011:LPK

[BS11b] Marcello Balduccini and Tran Cao Son, editors. Logic Programming, Knowledge Representation, and Nonmonotonic Reasoning: Essays Dedicated to Michael Gelfond on the Occasion of His 65th Birthday, volume 6565 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20831-2 (print), 3-642-20832-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20832-4.

Banerjee:2011:LAI

[BS11c] Mohua Banerjee and Anil Seth, editors. Logic and Its Applications: 4th Indian Conference, ICLA 2011, Delhi, India, January 5–11, 2011. Proceedings, volume 6521 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18025-6 (print), 3-642-18026-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18026-2.

Batyrshin:2011:AAI

[BS11d] Ildar Batyrshin and Grigori Sidorov, editors. Advances in Artificial Intelligence: 10th Mexican International Conference on Artificial Intelligence, MICAI 2011, Puebla, Mexico, November 26 — December 4, 2011, Proceedings, Part I, volume 7094 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25323-7 (print), 3-642-25324-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25324-9.

Batyrshin:2011:ASC

[BS11e] Ildar Batyrshin and Grigori Sidorov, editors. Advances in Soft Computing: 10th Mexican International Conference on Artificial Intelligence, MICAI 2011, Puebla, Mexico, November 26 — December 4, 2011, Proceedings, Part II, volume 7095 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25329-6 (print), 3-642-25330-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25330-0.

Butler:2011:FFM

[BS11f] Michael Butler and Wolfram Schulte, editors. FM 2011: Formal Methods: 17th International Symposium on Formal Methods, Limerick, Ireland, June 20–24, 2011. Proceedings, volume 6664 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21436-3 (print), 3-642-21437-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21437-0.

Bruni:2012:TGC

[BS12] Roberto Bruni and Vladimiro Sassone, editors. Trustworthy Global Computing: 6th International Symposium, TGC 2011, Aachen, Germany, June 9–10, 2011. Revised Selected Papers, volume 7173 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30064-2 (print), 3-642-30065-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30065-3.

Brueckner:2005:ESO

[BSA05] Sven A. Brueckner, Giovanna Di Marzo Serugendo, and Anthony, editors. Engineering self-organising systems: methodologies and applications, volume 3464 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26180-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26180-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3464;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136984.

Bouquet:1999:MUC

[BSB^{+}99] Paolo Bouquet, Luciano Serafini, Patrick Brézillon, Massimo Benerecetti, and Francesca Castellani, editors. Modeling and Using Context: Second International and Interdisciplinary Conference, CONTEXT’99, Trento, Italy, September 9–11, 1999: Proceedings, volume 1688 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66432-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1688. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1688.htm;
http://www.springerlink.
com/content/978-3-540-
66432-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1688.

Blache:2005:LAC

[BSBM05] Philippe Blache, Edward Stabler, Joan Busquets, and Richard Moot, editors. Logical aspects of computational linguistics: 5th International Conference, LACL 2005, Bordeaux, France, April 28–30, 2005: Proceedings, volume 3492 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25783-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN P98 .L3. URL http:

//www.springerlink.
com/content/978-3-540-
25783-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3492;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136076.

Balzarotti:2012:RAI

[BSC12] Davide Balzarotti, Salvatore J. Stolfo, and Marco Cova, editors. Research in Attacks, Intrusions, and Defenses: 15th International Symposium, RAID 2012, Amsterdam, The Netherlands, September 12–14, 2012. Proceedings, volume 7462 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33337-0 (print), 3-642-33338-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33338-5.

Bayer:1984:OSA

[BSG84] Rudolf Bayer, G. Seegmuller, and Robert M. Graham. Operating systems: an advanced course. Springer-Verlag Inc., New York, NY, USA, Springer study edition, 1984. CODEN LNCSD9. ISBN 3-540-09812-7. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 593 pp. LCCN QA76.6 .O6 1979. URL http:

//www.springerlink.
com/content/978-3-540-
09812-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3945.

Originally published as [GSB78].

Butz:2003:ABA

[BSG03] Martin V. Butz, Olivier Sigaud, and Pierre Gérard. Anticipatory Behavior in Adaptive Learning Systems: Foundations, Theories, and Systems, volume 2684 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40429-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 301 pp. LCCN Q325.5 .W65 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2684.htm;
 http://www.springerlink.
com/content/978-3-540-
40429-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2684.

From the preface: This book evolved out of the First Workshop on Adaptive Behavior in Anticipatory Learning Systems, held on the 11th of August in Edinburgh, UK, in conjunction with the 7th International Conference on Simulation of Adaptive Behavior: from Animals to Animats.

Brueckner:2006:ESO

[BSH06] Sven A. Brueckner, Giovanna Di Marzo Serugendo, and David Hales, editors. Engineering Self-Organising Systems: Third International Workshop, ESOA 2005, Utrecht, The Netherlands, July 25, 2005, Revised Selected Papers, volume 3910 of Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33342-8 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33342-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3910.

Bruza:2009:QIT

[BSL^{+}09] Peter Bruza, Donald Sofge, William Lawless, Keith van Rijsbergen, and Matthias Klusch, editors. Quantum Interaction: Third International Symposium, QI 2009, Saarbrücken, Germany, March 25–27, 2009. Proceedings, volume 5494 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00833-X (print), 3-642-00834-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00834-4.

Baar:2004:FUU

[BSMM04] Thomas Baar, Alfred Strohmeier, Ana Moreira, and Stephen J. Mellor, editors. ^{≪}UML^{≫}2004— The Unified Modelling Language: modelling languages and applications. 7th International Conference, Lisbon, Portugal, October 11–15, 2004. proceedings, volume 3273 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23307-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3273.htm;
 http://www.springerlink.
com/content/978-3-540-
23307-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3273;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101232.

Batten:2006:ISP

[BSN06] Lynn Margaret Batten and Reihaneh Safavi-Naini, editors. Information Security and Privacy: 11th Australasian Conference, ACISP 2006, Melbourne, Australia, July 3–5, 2006. Proceedings, volume 4058 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35458-1 (print), 3-540-35459-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35459-8.

Butz:2007:ABA

[BSPB07] Martin V. Butz, Olivier Sigaud, Giovanni Pezzulo, and Gianluca Baldassarre, editors. Anticipatory Behavior in Adaptive Learning Systems: From Brains to Individual and Social Behavior, volume 4520 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74261-1 (print), 3-540-74262-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74262-3.

Brusilovsky:2000:AHA

[BSS00] Peter Brusilovsky, Oliviero Stock, and Peter Strapparava, editors. Adaptive Hypermedia and Adaptive Web-Based Systems: International Conference, AH 2000, Trento, Italy, August 28–30, 2000. Proceedings, volume 1892 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67910-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I59 A36 2000.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1892.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67910-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1892.

Bjorner:2011:ADC

[BSS11] Nikolaj Bjørner and Viorica Sofronie-Stokkermans, editors. Automated Deduction — CADE-23: 23rd International Conference on Automated Deduction, Wrocław, Poland, July 31 — August 5, 2011. Proceedings, volume 6803 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22437-7 (print), 3-642-22438-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22438-6.

Balet:2001:VSU

[BST01] Olivier Balet, Gerard Subsol, and Patrice Torguet, editors. Virtual storytelling: using virtual reality technologies for storytelling: International Conference ICVS 2001, Avignon, France, September 27–28, 2001: Proceedings, volume 2197 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42611-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN PN212 .I68 2001; QA267.A1 L43 no.2197. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2197.htm;
 http://www.springerlink.
com/content/978-3-540-
42611-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2197.

Balet:2003:VSU

[BST03] Olivier Balet, Gérard Subsol, and Patrice Torguet, editors. Virtual Storytelling: Using Virtual Reality Techniques for Storytelling: Second International Conference, ICVS 2003, Toulouse, France, November 20–21, 2003: Proceedings, volume 2897 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20535-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN PN212 .I68 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2897.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20535-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2897;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b94275.

Berthold:2007:AID

[BSTL07] Michael R. Berthold, John Shawe-Taylor, and Nada Lavrač, editors. Advances in Intelligent Data Analysis VII: 7th International Symposium on Intelligent Data Analysis, IDA 2007, Ljubljana, Slovenia, September 6–8, 2007. Proceedings, volume 4723 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74824-5 (print), 3-540-74825-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74825-0.

Baldoni:2008:DAL

[BSvRW08] Matteo Baldoni, Tran Cao Son, M. Birna van Riemsdijk, and Michael Winikoff, editors. Declarative Agent Languages and Technologies V: 5th International Workshop, DALT 2007, Honolulu, HI, USA, May 14, 2007, Revised Selected and Invited Papers, volume 4897 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77563-3 (print), 3-540-77564-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77564-5.

Baldoni:2009:DAL

[BSvRW09] Matteo Baldoni, Tran Cao Son, M. Birna van Riemsdijk, and Michael Winikoff, editors. Declarative Agent Languages and Technologies VI: 6th International Workshop, DALT 2008, Estoril, Portugal, May 12, 2008, Revised Selected and Invited Papers, volume 5397 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-93919-9 (print), 3-540-93920-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
93920-7.

Barahona:1995:AIM

[BSW95a] P. Barahona, Mario Stefanelli, and J. Wyatt, editors. Artificial intelligence in medicine: 5th Conference on Artificial Intelligence in Medicine Europe, AIME ’95, Pavia, Italy, June 25–28, 1995: proceedings, volume 934 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60025-6, 0-387-60025-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R859.7.A78 C66 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0934.
htm;   http://www.springerlink.
com/content/978-3-540-
60025-1;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=934.

Bartosek:1995:STP

[BSW95b] Miroslav Bartosek, Jan Staudek, and J. Wiedermann, editors. SOF-SEM ’95: theory and practice of informatics: 22nd Seminar on Current Trends in Theory and Practice of Informatics, Milovy, Czech Republic, November 23– December 1, 1995: proceedings, volume 1012 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60609-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .S46 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1012.htm;
http://www.springerlink.
com/content/978-3-540-
60609-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1012.

Bubak:2012:BND

[BSW12] Marian Bubak, Tomasz Szepieniec, and Kazimierz Wiatr, editors. Building a National Distributed e-Infrastructure– PL-Grid: Scientific and Technical Achievements, volume 7136 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28266-0 (print), 3-642-28267-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28267-6.

Bao:2012:ACN

[BSZ12] Feng Bao, Pierangela Samarati, and Jianying Zhou, editors. Applied Cryptography and Network Security: 10th International Conference, ACNS 2012, Singapore, June 26–29, 2012. Proceedings, volume 7341 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31283-7 (print), 3-642-31284-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31284-7.

Burke:2006:PTA

[BT06] Edmund Burke and Michael Trick, editors. Practice and theory of automated timetabling V: 5th international conference, PATAT 2004, Pittsburgh, PA, USA, August 18–20, 2004, revised selected papers, volume 3616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30705-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30705-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3616.

Blanc-Talon:2008:ACI

[BTBP^{+}08] Jacques Blanc-Talon, Salah Bourennane, Wilfried Philips, Dan Popescu, and Paul Scheunders, editors. Advanced Concepts for Intelligent Vision Systems: 10th International Conference, ACIVS 2008, Juan-les-Pins, France, October 20–24, 2008. Proceedings, volume 5259 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88457-2 (print), 3-540-88458-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88458-3.

Blanc-Talon:2010:ACIa

[BTBP^{+}10a] Jacques Blanc-Talon, Don Bone, Wilfried Philips, Dan Popescu, and Paul Scheunders, editors. Advanced Concepts for Intelligent Vision Systems: 12th International Conference, ACIVS 2010, Sydney, Australia, December 13– 16, 2010, Proceedings, Part I, volume 6474 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17687-9 (print), 3-642-17688-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17688-3.

Blanc-Talon:2010:ACIb

[BTBP^{+}10b] Jacques Blanc-Talon, Don Bone, Wilfried Philips, Dan Popescu, and Paul Scheunders, editors. Advanced Concepts for Intelligent Vision Systems: 12th International Conference, ACIVS 2010, Sydney, Australia, December 13– 16, 2010, Proceedings, Part II, volume 6475 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17690-9 (print), 3-642-17691-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17691-3.

Bolc:2010:CVGa

[BTCW10a] Leonard Bolc, Ryszard Tadeusiewicz, Leszek J. Chmielewski, and Konrad Wojciechowski, editors. Computer Vision and Graphics: International Conference, ICCVG 2010, Warsaw, Poland, September 20– 22, 2010, Proceedings, Part I, volume 6374 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15909-5 (print), 3-642-15910-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15910-7.

Bolc:2010:CVGb

[BTCW10b] Leonard Bolc, Ryszard Tadeusiewicz, Leszek J. Chmielewski, and Konrad Wojciechowski, editors. Computer Vision and Graphics: International Conference, ICCVG 2010, Warsaw, Poland, September 20– 22, 2010, Proceedings, Part II, volume 6375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15906-0 (print), 3-642-15907-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15907-7.

Bolc:2012:CVG

[BTCW12] Leonard Bolc, Ryszard Tadeusiewicz, Leszek J. Chmielewski, and Konrad Wojciechowski, editors. Computer Vision and Graphics: International Conference, ICCVG 2012, Warsaw, Poland, September 24– 26, 2012. Proceedings, volume 7594 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33563-2 (print), 3-642-33564-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33564-8.

Bondavalli:2002:DCE

[BTF02] Andrea Bondavalli and Pascale Thevenod-Fosse, editors. Dependable computing EDCC-4: 4th European Dependable Computing Conference, Toulouse, France, October 23–25, 2002: Proceedings, volume 2485 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00012-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F38 E33 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2485.htm;
http://www.springerlink.
com/content/978-3-540-
00012-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2485.

Bussler:2005:SWD

[BTF05] Christoph Bussler, Val Tannen, and Irini Fundulaki, editors. Semantic Web and databases: second international workshop, SWDB 2004, Toronto, Canada, August 29–30, 2004. Revised selected papers, volume 3372 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24576-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815 .S94 2004. URL http:

//www.springerlink.
com/content/978-3-540-
24576-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3372;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106149.

Bruckstein:2012:SSV

[BtHRBB12] Alfred M. Bruckstein, Bart M. ter Haar Romeny, Alexander M. Bronstein, and Michael M. Bronstein, editors. Scale Space and Variational Methods in Computer Vision: Third International Conference, SSVM 2011, Ein-Gedi, Israel, May 29 — June 2, 2011, Revised Selected Papers, volume 6667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24784-9 (print), 3-642-24785-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24785-9.

Blanc-Talon:2011:ACI

[BTKP^{+}11] Jacques Blanc-Talon, Richard Kleihorst, Wilfried Philips, Dan Popescu, and Paul Scheunders, editors. Advances Concepts for Intelligent Vision Systems: 13th International Conference, ACIVS 2011, Ghent, Belgium, August 22– 25, 2011. Proceedings, volume 6915 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23686-3 (print), 3-642-23687-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23687-7.

Burmester:2011:ISI

[BTMI11] Mike Burmester, Gene Tsudik, Spyros Magliveras, and Ivana Ilić, editors. Information Security: 13th International Conference, ISC 2010, Boca Raton, FL, USA, October 25–28, 2010, Revised Selected Papers, volume 6531 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18177-5 (print), 3-642-18178-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18178-8.

Blanc-Talon:2012:ACI

[BTPP^{+}12] Jacques Blanc-Talon, Wilfried Philips, Dan Popescu, Paul Scheunders, and Pavel Zemčík, editors. Advanced Concepts for Intelligent Vision Systems: 14th International Conference, ACIVS 2012, Brno, Czech Republic, September 4–7, 2012. Proceedings, volume 7517 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33139-4 (print), 3-642-33140-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33140-4.

Blanc-Talon:2005:ACI

[BTPPS05] Jacques Blanc-Talon, Wilfried Philips, Dan Popescu, and Paul Scheunders, editors. Advanced concepts for intelligent vision systems: 7th international conference, ACIVS 2005, Antwerp, Belgium, September 20– 23, 2005: proceedings, volume 3708 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29032-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .A25 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29032-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3708.

Blanc-Talon:2006:ACI

[BTPPS06] Jacques Blanc-Talon, Wilfried Philips, Dan Popescu, and Paul Scheunders, editors. Advanced Concepts for Intelligent Vision Systems: 8th International Conference, ACIVS 2006, Antwerp, Belgium, September 18–21, 2006. Proceedings, volume 4179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44630-3 (print), 3-540-44632-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        44632-3.

Blanc-Talon:2007:ACI

[BTPPS07] Jacques Blanc-Talon, Wilfried Philips, Dan Popescu, and Paul Scheunders, editors. Advanced Concepts for Intelligent Vision Systems: 9th International Conference, ACIVS 2007, Delft, The Netherlands, August 28–31, 2007. Proceedings, volume 4678 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74606-4 (print), 3-540-74607-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74607-2.

Blanc-Talon:2009:ACI

[BTPPS09] Jacques Blanc-Talon, Wilfried Philips, Dan Popescu, and Paul Scheunders, editors. Advanced Concepts for Intelligent Vision Systems: 11th International Conference, ACIVS 2009, Bordeaux, France, September 28–October 2, 2009. Proceedings, volume 5807 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04696-7 (print), 3-642-04697-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04697-1.

Bauknecht:2002:CWT

[BTQ02] Kurt Bauknecht, A. Min Tjoa, and Gerald Quirchmayr, editors. Ecommerce and Web technologies: Third International Conference, EC-Web 2002, Aixen-Provence, France, September 2–6, 2002: Proceedings, volume 2455 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44137-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.32 .E35 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2455.htm;
 http://www.springerlink.
com/content/978-3-540-
44137-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2455.

Also available via the

World Wide Web.

Bauknecht:2003:CWT

[BTQ03] Kurt Bauknecht, A. Min Tjoa, and Gerald Quirchmayr, editors. E-Commerce and Web Technologies: 4th International Conference, EC-Web Prague, Czech Repbulic, September 2– 5, 2003 Proceedings, volume 2738 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40808-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.32 .I5796 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2738.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40808-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2738;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11826.

Bancilhon:1990:ADT

[BTT90] François Bancilhon, C. Thanos, and Dionysios C. Tsichritzis, editors. Advances in database technology — EDBT ’90: International Conference on Extending Database Technology, Venice, Italy, March 26–30, 1990, proceedings, volume 416 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-52291-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55844 1990.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0416.
htm;   http://www.springerlink.
com/content/978-0-387-
52291-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=416.

Brambilla:2012:WEI

[BTT12] Marco Brambilla, Takehiro Tokuda, and Robert Tolksdorf, editors. Web Engineering: 12th International Conference, ICWE 2012, Berlin, Germany, July 23–27, 2012. Proceedings, volume 7387 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31752-9 (print), 3-642-31753-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        31753-8.

Boll:2010:AMM

[BTZ^{+}10] Susanne Boll, Qi Tian, Lei Zhang, Zili Zhang, and Yi-Ping Phoebe Chen, editors. Advances in Multimedia Modeling: 16th International Multimedia Modeling Conference, MMM 2010, Chongqing, China, January 6–8, 2010. Proceedings, volume 5916 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11300-1 (print), 3-642-11301-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11301-7.

Bertino:1994:OOM

[BU94] Elisa Bertino and Susan Urban, editors. Object-oriented methodologies and systems: International Symposium, ISOOMS ’94, Palermo, Italy, September 21– 22, 1994: proceedings, volume 858 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58451-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 O24 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0858.htm;
 http://www.springerlink.
com/content/978-3-540-
58451-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=858.

Bohme:2001:IIC

[BU01] Thomas Böhme and Herwig Unger, editors. Innovative Internet computing systems: International Workshop IICS 2001, Ilmenau, Germany, June 21–22, 2001: Proceedings, volume 2060 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42275-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.59 .I59 2001; QA267.A1 L43 no.2060. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2060.htm;
 http://www.springerlink.
com/content/978-3-540-
42275-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2060.

Buchberger:1985:PEE

[Buc85] Bruno Buchberger, editor. Proceedings: EU-ROCAL ’85, European Conference on Computer Algebra, Linz, Austria, April 1–3, 1985. Volume 1. Invited Lectures, volume 203 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15983-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E85 1985.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0203.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        15983-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=203.

Buchmann:1990:DIL

[Buc90] A. Buchmann, editor. Design and implementation of large spatial databases: First Symposium SSD ’89, Santa Barbara, California, July 17–18, 1989: proceedings, volume 409 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52208-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26 S95 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0409.htm;
http://www.springerlink.
com/content/978-0-387-
52208-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=409.

Budach:1985:FCT

[Bud85] L. Budach, editor. Fundamentals of computation theory, volume 199 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CODEN LNCSD9. ISBN 0-387-15689-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I57 1985. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0199.
htm;   http://www.springerlink.
com/content/978-0-387-
15689-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=199.

Budach:1991:FCT

[Bud91] L. Budach, editor. Fundamentals of computation theory: 8th International Conference, FCT ’91, Gosen, Germany, September 9–13, 1991: proceedings, volume 529 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54458-5 (Berlin), 0-387-54458-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I573 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0529.htm;
 http://www.springerlink.
com/content/978-3-540-
54458-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=529.

Buell:2004:ANT

[Bue04] Duncan Buell, editor. Algorithmic Number Theory: 6th International Symposium, ANTS-VI, Burlington, VT, USA, June 13–18, 2004. Proceedings, volume 3076 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22156-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .H75 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3076.htm;

http://www.springerlink.
com/content/978-3-540-
22156-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3076;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98210.

Bass:1993:HCI

[BUG93] Leonard J. Bass, C. Unger, and Iu. M. Gornostaev, editors. Humancomputer interaction: third international conference, EWHCI ’93, Moscow, Russia, August 3–7, 1993: Selected Papers, volume 753 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57433-6 (Berlin), 0-387-57433-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I597 1993. DM72.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0753.
htm;   http://www.springerlink.
com/content/978-3-540-
57433-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=753.

Buhler:1998:ANT

[Buh98] Joe P. Buhler, editor.

Algorithmic number theory: Third International Symposium, ANTS-III, Portland, Oregon, USA, June 21–25, 1998: Proceedings, volume 1423 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64657-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L443 no.1423. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1423.htm;
 http://www.springerlink.
com/content/978-3-540-
64657-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1423.

Brinkschulte:2008:ACS

[BUHS08] Uwe Brinkschulte, Theo Ungerer, Christian Hochberger, and Rainer G. Spallek, editors. Architecture of Computing Systems

— ARCS 2008: 21st International Conference, Dresden, Germany, February 25–28, 2008. Proceedings, volume 4934 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78152-8 (print), 3-540-78153-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78153-0.

Bui:1987:COG

[Bui87] Tung X. Bui. Co-oP:

a group decision support system for cooperative multiple criteria group decision making, volume 290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-18753-7 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 250 pp. LCCN T58.62 .B851 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0290.htm;
http://www.springerlink.
com/content/978-0-387-
18753-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=290.

Bundy:1994:ADC

[Bun94] Alan Bundy, editor.

Automated deduction, CADE-12: 12th International Conference on Automated Deduction, Nancy, France, June 26–July 1, 1994: proceedings, volume 814 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58156-1 (Berlin), 0-387-58156-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D337 I58 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0814.htm;
 http://www.springerlink.
com/content/978-3-540-
58156-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=814.

Buning:1996:CSL

[Bun96] H. (Hans) Kleine Buning, editor. Computer science logic: 9th International Workshop, CSL ’95, Annual Conference of the EACSL, Paderborn, Germany, September 22–29, 1995: Selected Papers, volume 1092 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61377-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W625 1995. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t1092.htm;
http://www.springerlink.
com/content/978-3-540-
61377-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1092.

Burkhart:1990:CVI

[Bur90] H. Burkhart, editor.

CONPAR 90-VAPP IV: Joint International Conference on Vector and Parallel Processing, Zürich, Switzerland, September 10–13, 1990: proceedings, volume 457 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-53065-7 (Berlin), 0-387-53065-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .J65 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0457.htm;
http://www.springerlink.
com/content/978-3-540-
53065-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=457.

Burckert:1991:RPL

[Bur91] H. J. (Hans-Jurgen) Burckert. A resolution principle for a logic with restricted quantifiers, volume 568 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-55034-8 (Berlin), 0-387-55034-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 116 pp. LCCN QA76.63 .B8 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0568.htm;
 http://www.springerlink.
com/content/978-3-540-
55034-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=568.

Burkart:1997:AVS

[Bur97] Olaf Burkart. Automatic verification of sequential infinite-state processes, volume 1354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63982-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 163 pp. LCCN QA76.76.V47 B87 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1354.htm;
 http://www.springerlink.

com/content/978-3-540-
63982-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1354.

Businger:1985:PLD

[Bus85] Arnold Businger. POR-TAL language description, volume 198 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15682-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 186 pp. LCCN QA76.73.P66 B871 1985. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0198.htm;
 http://www.springerlink.
com/content/978-0-387-
15682-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=198.

Businger:1988:PLD

[Bus88] Arnold Businger. POR-TAL language description, volume 198 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, second extended edition, 1988. CO-DEN LNCSD9. ISBN 3-540-18960-2 (Berlin), 0-387-18960-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 197 pp. LCCN QA267.A1 L43 no.198 suppl. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0198.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        18960-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=198.

Butler:1991:FAP

[But91] G. (Gregory) Butler. Fundamental algorithms for permutation groups, volume 559 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54955-2 (Berlin), 0-387-54955-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 238 pp. LCCN QA175 .B88 1991. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0559.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        54955-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=559.

Butterfield:2010:UTP

[But10] Andrew Butterfield, editor. Unifying Theories of Programming: Second International Symposium, UTP 2008, Dublin, Ireland, September 8–10, 2008, Revised Selected Papers, volume 5713 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14520-5 (print), 3-642-14521-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14521-6.

Bensoussan:1992:FTC

[BV92] Alain Bensoussan and J.- P Verjus, editors. Future tendencies in computer science, control, and applied mathematics: International Conference on the Occasion of the 25th Anniversary of IN-RIA: Paris, France, December 8–11, 1992: proceedings, volume 653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-56320-2, 0-387-56320-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I5484 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0653.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        56320-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=653.

Buchberger:1994:PPC

[BV94] Bruno Buchberger and Jens Volkert, editors. Parallel processing: CON-PAR 94-VAPP VI: third Joint International Conference on Vector and Parallel Processing, Linz, Austria, September 6– 8, 1994: proceedings, volume 854 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58430-7 (Berlin), 0-387-58430-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .J65 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0854.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58430-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=854.

Boman:1997:MAR

[Bv97] Magnus Boman and Walter van de Velde, editors. Multi-agent rationality: 8th European Workshop on Modelling Autonomous Agents in a Multi-Agent World, MAAMAW’97, Ronneby, Sweden, May 13–16, 1997: proceedings, volume 1237 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63077-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 E87 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1237.htm;
 http://www.springerlink.
com/content/978-3-540-
63077-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1237.

Baaz:2002:LPA

[BV02] Matthias Baaz and A. Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 9th International Conference, LPAR 2002, Tbilisi, Georgia, October 14–18, 2002. Proceedings, volume 2514 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00010-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I55 2002; QA76.63 .L73 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2514.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00010-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2514.

Baader:2005:LPA

[BV05a] Franz Baader and Andrei Voronkov, editors. Logic for programming, artificial intelligence, and reasoning: 11th international conference, LPAR 2004, Montevideo, Uruguay, March 14–18, 2005, proceedings, volume 3452 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25236-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .L735.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        25236-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3452;

http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106931.

Brun:2005:GBR

[BV05b] Luc Brun and Mario Vento, editors. Graphbased representations in pattern recognition: 5th IAPR international workshop, GbRPR 2005, Poitiers, France, April 11–13, 2005: Proceedings, volume 3434 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25270-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1650 .G37 2005. URL http:

//www.springerlink.
com/content/978-3-540-
25270-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3434;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107037.

Bishop:2011:OMC

[BV11] Judith Bishop and Antonio Vallecillo, editors. Objects, Models, Components, Patterns: 49th International Conference, TOOLS 2011, Zurich, Switzerland, June 28– 30, 2011. Proceedings, volume 6705 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21951-9 (print), 3-642-21952-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21952-8.

Bjorner:2012:LPA

[BV12] Nikolaj Bjørner and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 18th International Conference, LPAR-18, Mérida, Venezuela, March 11– 15, 2012. Proceedings, volume 7180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28716-6 (print), 3-642-28717-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28717-6.

Bubak:2008:CSIa

[BvADS08a] Marian Bubak, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science — ICCS 2008: 8th International Conference, Kraków, Poland, June 23–25, 2008, Proceedings, Part I, volume 5101 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69383-1 (print), 3-540-69384-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69384-0.

Bubak:2008:CSIb

[BvADS08b] Marian Bubak, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science — ICCS 2008: 8th International Conference, Kraków, Poland, June 23–25, 2008, Proceedings, Part II, volume 5102 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69386-6 (print), 3-540-69387-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        69387-1.

Bubak:2008:CSI

[BvADS08c] Marian Bubak, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational science — ICCS 2008. 8th international conference, Kraków, Poland, June 23–25, 2008. Proceedings, Part III, volume 5103 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008.

Bubak:2004:CSIa

[BvASD04a] Marian Bubak, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational Science—ICCS 2004: 4th International Conference Kraków, Poland, June 6–9, 2004 Proceedings, Part I, volume 3036 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22114-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3036.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-

22114-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3036;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97987.

Bubak:2004:CSIb

[BvASD04b] Marian Bubak, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational Science—ICCS 2004: 4th International Conference Kraków, Poland, June 6–9, 2004 Proceedings, Part II, volume 3037 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22115-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3037.htm;
 http://www.springerlink.
com/content/978-3-540-
22115-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3037;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97988.

Bubak:2004:CSIc

[BvASD04c] Marian Bubak, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational Science—ICCS 2004: 4th International Conference Kraków, Poland, June 6–9, 2004 Proceedings, Part III, volume 3038 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22116-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3038.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22116-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3038;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97989.

Bubak:2004:CSId

[BvASD04d] Marian Bubak, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational Science—ICCS 2004: 4th International Conference Kraków, Poland, June 6–9, 2004 Proceedings, Part IV, volume 3039 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22129-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3039.htm;
 http://www.springerlink.
com/content/978-3-540-
22129-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3039;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98005.

Biganzoli:2012:CIM

[BVAT12] Elia Biganzoli, Alfredo Vellido, Federico Ambrogi, and Roberto Tagliaferri, editors. Computational Intelligence Methods for Bioinformatics and Biostatistics: 8th International Meeting, CIBB 2011, Gargnano del Garda, Italy, June 30 — July 2, 2011, Revised Selected Papers, volume 7548 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35685-0 (print), 3-642-35686-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        35686-5.

Braubach:2009:MST

[BvdHPP09] Lars Braubach, Wiebe van der Hoek, Paolo Petta, and Alexander Pokahr, editors. Multiagent System Technologies: 7th German Conference, MATES 2009, Hamburg, Germany, September 9– 11, 2009. Proceedings, volume 5774 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04142-6 (print), 3-642-04143-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04143-3.

Blackburn:2011:TTL

[BvDMST11] Patrick Blackburn, Hans van Ditmarsch, María Manzano, and Fernando Soler-Toscano, editors. Tools for Teaching Logic: Third International Congress, TICTTL 2011, Salamanca, Spain, June 1– 4, 2011. Proceedings, volume 6680 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21349-9 (print), 3-642-21350-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21350-2.

Bousquet:2004:ALM

[BvLR04] Olivier Bousquet, Ulrike von Luxburg, and Gunnar Rätsch, editors. Advanced Lectures on Machine Learning: ML summer schools 2003, Canberra, Australia, February 2–14, 2003, Tuebingen, Germany, August 4–16, 2003, Revised Lectures, volume 3176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23122-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3176.htm;
 http://www.springerlink.
com/content/978-3-540-
23122-6;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3176;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100712.

Brandenburg:1987:SAS

[BVNW87] F. J. Brandenburg, G. Vidal-Naquet, and M. Wirsing, editors. STACS 87: 4th Annual Symposium on Theoretical Aspects of Computer Science, Passau, Federal Republic of Germany, February 19–21, 1987, proceedings, volume 247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-17219-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.247. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0247.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        17219-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=247.

Babar:2010:PFS

[BVO10] M. Ali Babar, Matias Vierimaa, and Markku Oivo, editors. Product-Focused Software Process Improvement: 11th International Conference, PROFES 2010, Limerick, Ireland, June 21– 23, 2010. Proceedings, volume 6156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13791-1 (print), 3-642-13792-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13792-1.

Broy:1991:MPS

[BW91] M. Broy and M. Wirsing.

Methods of programming: selected papers on the CIP-project, volume 544 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54576-X (Berlin), 0-387-54576-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 268 pp. LCCN QA76.6 .M4543 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0544.htm;
 http://www.springerlink.
com/content/978-3-540-
54576-7;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=544.

Bruynooghe:1992:PLI

[BW92] M. Bruynooghe and M. Wirsing, editors. Programming language implementation and logic programming: 4th International Symposium, PLILP ’92, Leuven, Belgium, August 26– 28, 1992: proceedings, volume 631 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 0-387-55844-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .I5773 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0631.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        55844-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=631.

Burckard:1997:AEA

[BW97] Rainer E. Burckard and Gerhard Woeginger, editors. Algorithms—ESA ’97: 5th annual European symposium, Graz, Austria, September 15– 17, 1997: proceedings, volume 1284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63397-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1284.htm;
 http://www.springerlink.
com/content/978-3-540-
63397-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1284.

Brandes:2000:GTC

[BW00] Ulrik Brandes and Dorothea Wagner, editors. Graph-Theoretic Concepts in Computer Science: 26th International Workshop, WG 2000, Konstanz, Germany, June 15–17, 2000. Proceedings, volume 1928 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41183-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .I572 2000; QA75.5 .I647 2000; QA267.A1 L43 no.1928. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1928.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        41183-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1928.

Brebner:2001:FPL

[BW01] Gordon Brebner and Roger Woods, editors. Field-programmable logic and applications: 11th International Conference, FPL 2001, Belfast, Northern Ireland, UK, August 27–29, 2001: Proceedings, volume 2147 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42499-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.G36 I48 2001; QA267.A1 L43 no.2147. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2147.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42499-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2147.

Bai:2003:EIC

[BW03a] Fengshan Bai and Bernd Wegner, editors. Electronic Information and Communication in Mathematics: ICM 2002 International Satellite Conference, Beijing, China, August 29–31, 2002: Revised Papers, volume 2730 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40689-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA41.5 .E54 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2730.htm;
 http://www.springerlink.
com/content/978-3-540-
40689-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2730;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11818.

Basin:2003:TPH

[BW03b] David Basin and Burkhart Wolff, editors. Theorem Proving in Higher Order Logics: 16th International Conference, TPHOLs 2003, Rome, Italy, September 8–12, 2003: Proceedings, volume 2758 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40664-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2758.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40664-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2758;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11828.

Bacchus:2005:TAS

[BW05a] Fahiem Bacchus and Toby Walsh, editors. Theory and Applications of Satisfiability Testing: 8th International Conference, SAT 2005, St Andrews, UK, June 19– 23, 2005. Proceedings, volume 3569 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26276-8 (print), 3-540-31679-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3569.

Briand:2005:MDE

[BW05b] Lionel Briand and Clay Williams, editors. Model driven engineering languages and systems: 8th international conference, Models 2005, Montego Bay, Jamaica, October 2–7, 2005, proceedings, volume 3713 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29010-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29010-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3713.

Bruni:2009:WSF

[BW09] Roberto Bruni and Karsten Wolf, editors. Web Services and Formal Methods: 5th International Workshop, WS-FM 2008, Milan, Italy, September 4–5, 2008, Revised Selected Papers, volume 5387 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01363-5 (print), 3-642-01364-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        01364-5.

Baudry:2010:SCI

[BW10a] Benoît Baudry and Eric Wohlstadter, editors. Software Composition: 9th International Conference, SC 2010, Malaga, Spain, July 1–2, 2010. Proceedings, volume 6144 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14045-9 (print), 3-642-14046-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14046-4.

Bi:2010:KSE

[BW10b] Yaxin Bi and Mary-Anne Williams, editors. Knowledge Science, Engineering and Management: 4th International Conference, KSEM 2010, Belfast, Northern Ireland, UK, September 1– 3, 2010. Proceedings, volume 6291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15279-1 (print), 3-642-15280-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15280-1.

Bao:2011:ISP

[BW11] Feng Bao and Jian Weng, editors. Information Security Practice and Experience: 7th International Conference, IS-PEC 2011, Guangzhou, China, May 30 — June 1, 2011. Proceedings, volume 6672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21030-9 (print), 3-642-21031-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21031-0.

Becker:2009:RCA

[BWAM09] Jürgen Becker, Roger Woods, Peter Athanas, and Fearghal Morgan, editors. Reconfigurable Computing: Architectures, Tools and Applications: 5th International Workshop, ARC 2009, Karlsruhe, Germany, March 16–18, 2009. Proceedings, volume 5453 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00640-X (print), 3-642-00641-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00641-8.

Braspenning:1995:ANN

[BWT95] P. J. Braspenning, A. J. M. M. Weijters, and F. Thuijsman. Artificial neural networks: an introduction to ANN theory and practice, volume 931 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59488-4 (Berlin), 0-387-59488-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 293 pp. LCCN QA76.87 .A736 1995.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0931.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        59488-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=931.

Burmester:2006:SMA

[BY06] Mike Burmester and Alec Yasinsac, editors. Secure Mobile Ad-hoc Networks and Sensors: First International Workshop, MADNES 2005, Singapore, September 20–22, 2005, Revised Selected Papers, volume 4074 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36646-6 (print), 3-540-37863-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37863-1.

Biham:2007:SAC

[BY07] Eli Biham and Amr M.

Youssef, editors. Selected Areas in Cryptography: 13th International Workshop, SAC 2006, Montreal, Canada, August 17–18, 2006 Revised Selected Papers, volume 4356 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74461-4 (print), 3-540-74462-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74462-7.

Baeza-Yates:2003:CPM

[BYCC03] R. Baeza-Yates, Edgar Chávez, and Maxime Crochemore, editors. Combinatorial pattern matching: 14th annual symposium, CPM 2003, Morelia, Michoacan, Mexico, June 25– 27, 2003: Proceedings, volume 2676 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40311-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2676.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40311-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2676.

Also available via the

World Wide Web.

Baeza-Yates:2012:AIR

[BYdVZ^{+}12] Ricardo Baeza-Yates, Arjen P. de Vries, Hugo Zaragoza, B. Barla Cambazoglu, Vanessa Murdock, Ronny Lempel, and Fabrizio Silvestri, editors. Advances in Information Retrieval: 34th European Conference on IR Research, ECIR 2012, Barcelona, Spain, April 1–5, 2012. Proceedings, volume 7224 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28996-7 (print), 3-642-28997-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28997-2.

Baeza-Yates:1995:LTI

[BYGP95] R. Baeza-Yates, E. Goles, and Patricio V. Poblete, editors. LATIN ’95: theoretical informatics: second Latin American Symposium, Valparaiso, Chile, April 3–7, 1995: proceedings, volume 911 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59175-3 (Berlin), 0-387-59175-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .L348 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0911.htm;
 http://www.springerlink.
com/content/978-3-540-
59175-7;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=911.

Bao:2011:ISC

[BYLJ11] Feng Bao, Moti Yung, Dongdai Lin, and Jiwu Jing, editors. Information Security and Cryptology: 5th International Conference, Inscrypt 2009, Beijing, China, December 12–15, 2009. Revised Selected Papers, volume 6151 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16341-6 (print), 3-642-16342-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16342-5.

Blaser:1983:EST

[BZ83] A. Blaser and M. Zoeppritz, editors. Enduser systems and their human factors: proceedings of the scientific symposium conducted on the occasion of the 15th anniversary of the Science Center Heidelberg of IBM Germany, Heidelberg, March 18, 1983, volume 150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-12273-7 (paperback), 3-540-12273-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.I58 E53 1983. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0150.htm;
 http://www.springerlink.
com/content/978-0-387-
12273-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=150.

Broy:2003:PSI

[BZ03] Manfred Broy and Alexandre V. Zamulin, editors. Perspectives of System Informatics: 5th International Andrei Ershov Memorial Conference, PSI 2003, Akademgorodok, Novosibirsk, Russia, July 9– 12, 2003: Revised Papers, volume 2890 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20813-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I5443 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2890.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        20813-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2890;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b94823.

Bourgeois:2008:AAP

[BZ08a] Anu G. Bourgeois and S. Q. Zheng, editors. Algorithms and Architectures for Parallel Processing: 8th International Conference, ICA3PP 2008, Cyprus, June 9–11, 2008 Proceedings, volume 5022 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69500-1 (print), 3-540-69501-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69501-1.

Buning:2008:TAS

[BZ08b] Hans Kleine Büning and Xishun Zhao, editors. Theory and Applications of Satisfiability Testing

– SAT 2008: 11th International Conference, SAT 2008, Guangzhou, China, May 12–15, 2008. Proceedings, volume 4996 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79718-1 (print), 3-540-79719-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79719-7.

Bravetti:2009:CCT

[BZ09] Mario Bravetti and Gianluigi Zavattaro, editors. CONCUR 2009 — Concurrency Theory: 20th International Conference, CONCUR 2009, Bologna, Italy, September 1–4, 2009. Proceedings, volume 5710 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04080-2 (print), 3-642-04081-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04081-8.

Bin:2007:HSV

[BZU07] Eyal Bin, Avi Ziv, and Shmuel Ur, editors. Hardware and Software, Verification and Testing: Second International Haifa Verification Conference, HVC 2006, Haifa, Israel, October 23–26, 2006. Revised Selected Papers, volume 4383 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70888-X (print), 3-540-70889-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70889-6.

Cai:2006:AIE

[CA06] Yang Cai and Julio Abascal, editors. Ambient Intelligence in Everyday Life: Foreword by Emile Aarts, volume 3864 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37785-9 (print), 3-540-37788-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37788-7.

Cerda-Alabern:2008:WSM

[CA08] Llorenç Cerdà-Alabern, editor. Wireless Systems and Mobility in Next Generation Internet: 4th International Workshop of the EuroNGI/EuroFGI Network of Excellence Barcelona, Spain, January 16–18, 2008 Revised Selected Papers, volume 5122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89182-X (print), 3-540-89183-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89183-3.

Clarke:2010:CML

[CA10] Dave Clarke and Gul Agha, editors. Coordination Models and Languages: 12th International Conference, COORDINATION 2010, Amsterdam, The Netherlands, June 7–9, 2010. Proceedings, volume 6116 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13413-0 (print), 3-642-13414-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        13414-2.

Cohen:2010:AID

[CAB10] Paul R. Cohen, Niall M. Adams, and Michael R. Berthold, editors. Advances in Intelligent Data Analysis IX: 9th International Symposium, IDA 2010, Tucson, AZ, USA, May 19–21, 2010. Proceedings, volume 6065 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13061-5 (print), 3-642-13062-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13062-5.

Cadoli:1995:TRA

[Cad95] Marco Cadoli. Tractable reasoning in artificial intelligence, volume 941 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60058-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 247 pp. LCCN Q335 .C23 1995. URL

                                        http://link.springer-

ny.com/link/service/
series/0558/tocs/t0941.
htm;   http://www.springerlink.
com/content/978-3-540-
60058-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=941.

Caelli:2002:SSS

[CAD^{+}02] Terry Caelli, Adnan Amin, Robert P. W. Duin, Mohamed Kamel, and Dick de Ridder, editors. Structural, Syntactic, and Statistical Pattern Recognition: Joint IAPR International Workshops SSPR 2002 and SPR 2002, Windsor, Ontario, Canada, August 6–9, 2002. Proceedings, volume 2396 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44011-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 .I57 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2396.htm;
 http://www.springerlink.
com/content/978-3-540-
44011-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2396.

Cordella:2001:VFI

[CAdB01] Luigi P. Cordella, Carlo Arcelli, and Gabriella Sanniti di Baja, editors. Visual form 2001: 4th International Workshop on Visual Form, IWVF4, Capri, Italy, May 28– 30, 2001: Proceedings, volume 2059 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42120-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .I56 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2059.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42120-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2059.

Cagnoni:2000:RWA

[Cag00] Stefano Cagnoni, editor. Real-world applications of evolutionary computing: EvoWorkshops 2000: EvoIASP, EvoSCONDI, EvoTel, EvoSTIM, EvoRob, and EvoFlight, Edinburgh, Scotland, UK, April 17, 2000: Proceedings, volume 1803 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67353-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E899 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1803.htm;
 http://www.springerlink.
com/content/978-3-540-
67353-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1803.

Cagnoni:2002:AEC

[Cag02] Stefano Cagnoni, editor. Applications of evolutionary computing: EvoWorkshops 2002: Evo-COP, EvoIASP, EvoS-TIM/EvoPLAN, Kinsale, Ireland, April 3– 4, 2002: Proceedings, volume 2279 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43432-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E899 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2279.htm;
 http://www.springerlink.
com/content/978-3-540-
43432-0;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2279.

Held in conjunction with the 5th European Conference on Genetic Programming.

Cai:2005:AIS

[Cai05] Yang Cai, editor. Ambient Intelligence for Scientific Discovery: Foundations, Theories, and Systems, volume 3345 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24466-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5915 .S54 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24466-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3345;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105582.

Cai:2008:DHM

[Cai08] Yang Cai, editor. Digital Human Modeling: Trends in Human Algorithms, volume 4650 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89429-2 (print), 3-540-89430-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89430-8.

Cai:2011:CIR

[Cai11] Yang Cai, editor. Computing with Instinct: Rediscovering Artificial Intelligence, volume 5897 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19756-6 (print), 3-642-19757-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19757-4.

Calmet:1982:CAE

[Cal82] Jacques Calmet, editor. Computer algebra: EUROCAM’82, European Computer Algebra Conference, Marseille, France, 5–7 April, 1982, volume 144 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11607-9 (USA).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.144. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0144.htm;
 http://www.springerlink.
com/content/978-0-387-
11607-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=144.

Calmet:1986:AAE

[Cal86] Jacques Calmet, editor. Algebraic algorithms and error-correcting codes: 3rd international conference, AAECC-3, Grenoble, France, July 15– 19, 1985, proceedings, volume 229 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16776-5 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 A551 1986.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0229.
htm;   http://www.springerlink.
com/content/978-0-387-
16776-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=229.

“Organized by the ‘Laboratoire d’informatique fondamentale et d’intelligence artificielle’ (LIFIA)”– Pref.

Cantoni:1989:RIP

[Can89] V. Cantoni, editor. Recent issues in pattern analysis and recognition, volume 399 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51815-0 (New York), 3-540-51815-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 R433 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0399.htm;
http://www.springerlink.
com/content/978-0-387-
51815-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=399.

Papers from a satellite conference held in Rome, Nov. 14–17, 1988, and from work done under the patronage of the Academy of Science, Berlin, German Democratic Republic.

Canetti:2008:TCF

[Can08] Ran Canetti, editor.

Theory of Cryptography: Fifth Theory of Cryptography Conference, TCC 2008, New York, USA, March 19–21, 2008. Proceedings, volume 4948 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78523-X (print), 3-540-78524-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78524-8.

Canteaut:2012:FSE

[Can12] Anne Canteaut, editor. Fast Software Encryption: 19th International Workshop, FSE 2012, Washington, DC, USA, March 19–21, 2012. Revised Selected Papers, volume 7549 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34046-6 (print), 3-642-34047-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34047-5.

Corchado:2008:HAI

[CAP08] Emilio Corchado, Ajith Abraham, and Witold Pedrycz, editors. Hybrid Artificial Intelligence Systems: Third International Workshop, HAIS 2008, Burgos, Spain, September 24– 26, 2008. Proceedings, volume 5271 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87655-3 (print), 3-540-87656-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87656-4.

Cardelli:2003:EOO

[Car03] Luca Cardelli, editor.

ECOOP 2003 — Objectoriented programming: 17th European Conference, Darmstadt, Germany, July 21–25, 2003: Proceedings, volume 2743 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40531-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35 E33 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2743.
htm;   http://www.springerlink.
com/content/978-3-540-
40531-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2742.

Casanova:1981:CCP

[Cas81] Marco Antonio Casanova. The concurrency control problem for database systems, volume 116 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10845-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 175 pp. LCCN QA76.9.D3 .C37. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0116.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        10845-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=116.

Cassez:2001:MVP

[Cas01] Franck Cassez, editor. Modeling and verification of parallel processes: 4th Summer School, MOVEP 2000, Nantes, France, June 19–23, 2000, volume 2067 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42787-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .M69 2000; QA267.A1 L43 no.2067. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2067.htm;
http://www.springerlink.
com/content/978-3-540-
42787-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2067.

Castagna:2009:PLS

[Cas09a] Giuseppe Castagna, editor. Programming Languages and Systems: 18th European Symposium on Programming, ESOP 2009, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2009, York, UK, March 22–29, 2009. Proceedings, volume 5502 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00589-6 (print), 3-642-00590-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00590-9.

Cesar:2009:EPW

[CAS^{+}09b] Eduardo César, Michael Alexander, Achim Streit, Jesper Larsson Träff, Christophe Cérin, Andreas Knüpfer, Dieter Kranzlmüller, and Shantenu Jha, editors. Euro-Par 2008 Workshops — Parallel Processing: VHPC 2008, UNICORE 2008, HPPC 2008, SGS 2008, PROPER 2008, ROIA 2008, and DPA 2008, Las Palmas de Gran Canaria, Spain, August 25–26, 2008, Revised Selected Papers, volume 5415 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00954-9 (print), 3-642-00955-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00955-6.

Caviness:1985:PEE

[Cav85] Bob F. Caviness, editor. Proceedings: EU-ROCAL ’85, European Conference on Computer Algebra, Linz, Austria, April 1–3, 1985. Volume 2. Research Contributions, volume 204 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15984-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E85 1985.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0204.
htm;   http://www.springerlink.
com/content/978-0-387-
15984-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=204.

CoelloCoello:2005:EMC

[CAZ05] Carlos A. Coello Coello, Arturo Hernández Aguirre, and Eckart Zitzler, editors. Evolutionary multi-criterion optimization: third international conference, EMO 2005, Guanajuato, Mexico, March 9–11, 2005. Proceedings, volume 3410 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24983-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T57.95 .E48 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24983-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3410;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106458.

Cimatti:2006:FMH

[CB06] Alessandro Cimatti and Marco Bernardo, editors. Formal methods for hardware verification: 6th International School on Formal Methods for the Design of Computer, Communication, and Software Systems, SFM 2006, Bertinoro, Italy, May 22– 27, 2006: advanced lectures, volume 3965 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34304-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 I586 2006.

                                        URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0661/2006925529-d.
                                        html;   http://www.
                                        loc.gov/catdir/toc/
                                        fy0705/2006925529.html;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        34304-2.

Ceri:2010:SCC

[CB10] Stefano Ceri and Marco Brambilla, editors. Search Computing: Challenges and Directions, volume 5950 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12309-0 (print), 3-642-12310-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12310-8.

Ceri:2011:SCT

[CB11] Stefano Ceri and Marco Brambilla, editors. Search Computing: Trends and Developments, volume 6585 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19667-5 (print), 3-642-19668-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19668-3.

Ceri:2012:SCB

[CB12] Stefano Ceri and Marco Brambilla, editors. Search Computing: Broadening Web Search, volume 7538 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34212-4 (print), 3-642-34213-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        34213-4.

Cremers:2009:EMM

[CBBS09] Daniel Cremers, Yuri Boykov, Andrew Blake, and Frank R. Schmidt, editors. Energy Minimization Methods in Computer Vision and Pattern Recognition: 7th International Conference, EMMCVPR 2009, Bonn, Germany, August 24–27, 2009. Proceedings, volume 5681 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03640-6 (print), 3-642-03641-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03641-5.

Calmet:2002:AIA

[CBC^{+}02] Jacques Calmet, Belaid Benhamou, Olga Caprotti, Laurent Henocque, and Volker Sorge, editors. Artificial Intelligence, Automated Reasoning, and Symbolic Computation: Joint International Conference, AISC 2002 and Calculemus 2002 Marseille, France, July 1–5, 2002. Proceedings, volume 2385 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43865-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A768 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2385.htm;
 http://www.springerlink.
com/content/978-3-540-
43865-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2385.

Cuppens-Boulahia:2012:DAS

[CBCGA12] Nora Cuppens-Boulahia, Frédéric Cuppens, and Joaquin Garcia-Alfaro, editors. Data and Applications Security and Privacy XXVI: 26th Annual IFIP WG 11.3 Conference, DBSec 2012, Paris, France, July 11–13,2012. Proceedings, volume 7371 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31539-9 (print), 3-642-31540-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        31540-4.

Comon:2001:CAV

[CBF01] Hubert Comon, Gerard Berry, and A. Finkel, editors. Computer Aided Verification: 13th International Conference, CAV 2001, Paris, France, July 18–22, 2001: Proceedings, volume 2102 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42345-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E93 C38 2001.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2102.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        42345-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2102.

Carrico:2009:GDI

[CBF09] Luís Carriço, Nelson Baloian, and Benjamim Fonseca, editors. Groupware: Design, Implementation, and Use: 15th International Workshop, CRIWG 2009, Peso da Régua, Douro, Portugal, September 13–17, 2009. Proceedings, volume 5784 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04215-5 (print), 3-642-04216-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04216-4.

Cao:2010:ADMa

[CBG^{+}10] Longbing Cao, Ana L. C. Bazzan, Vladimir Gorodetsky, Pericles A. Mitkas, Gerhard Weiss, and Philip S. Yu, editors. Agents and Data Mining Interaction: 6th International Workshop on Agents and Data Mining Interaction, ADMI 2010, Toronto, ON, Canada, May 11, 2010, Revised Selected Papers, volume 5980 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15419-0 (print), 3-642-15420-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15420-1.

Calderon-Benavides:2012:SPI

[CBGCCZ12] Liliana Calderón-Benavides, Cristina González-Caro, Edgar Chávez, and Nivio Ziviani, editors. String Processing and Information Retrieval: 19th International Symposium, SPIRE 2012, Cartagena de Indias, Colombia, October 21–25, 2012. Proceedings, volume 7608 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34108-X (print), 3-642-34109-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34109-0.

Christensen:1999:SBI

[CBN99] H. I. Christensen, Horst Bunke, and Hartmut Noltemeier, editors. Sensor based intelligent robots: International Workshop, Dagstuhl Castle, Germany, September 28–October 2, 1998: Selected Papers, volume 1724 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66933-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN TJ211 .S437 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1724.htm;
 http://www.springerlink.
com/content/978-3-540-
66933-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1724.

Cesa-Bianchi:2002:FLT

[CBNR02] Nicolo Cesa-Bianchi, Masayuki Numao, and Rudiger Reischuk, editors. FAlgorithmic Learning Theory: 13th International Conference, ALT 2002, Lübeck, Germany, November 24–26, 2002. Proceedings, volume 2533 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00170-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A48 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2533.
htm;   http://www.springerlink.
com/content/978-3-540-
00170-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2533.

Also available via the World Wide Web.

Charron-Bost:2010:RTP

[CBPS10]  Bernadette  Charron-

Bost, Fernando Pedone, and André Schiper, editors. Replication: Theory and Practice, volume 5959 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11293-5 (print), 3-642-11294-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11294-2.

Coden:2002:IRT

[CBS02] Anni R. Coden, Eric W. Brown, and Savitha Srinivasan, editors. Information Retrieval Techniques for Speech Applications, volume 2273 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43156-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 I54 2002; TK7882.S65 I54 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2273.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-

43156-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2273.

Based on a workshop held on September 13, 2001 held as part of the 24th Annual International ACM SIGIR Conference on Research and Development in Information Retrieval.

Chiu:2011:WIS

[CBS^{+}11] Dickson K. W. Chiu, Ladjel Bellatreche, Hideyasu Sasaki, Ho fung Leung, Shing-Chi Cheung, Haiyang Hu, and Jie Shao, editors. Web Information Systems Engineering — WISE 2010 Workshops: WISE 2010 International Symposium WISS, and International Workshops CISE, MBC, Hong Kong, China, December 12–14, 2010, Revised Selected Papers, volume 6724 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24395-9 (print), 3-642-24396-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24396-7.

Cao:2012:ADM

[CBS^{+}12] Longbing Cao, Ana L. C. Bazzan, Andreas L. Symeonidis, Vladimir I. Gorodetsky, Gerhard Weiss, and Philip S. Yu, editors. Agents and Data Mining Interaction: 7th International Workshop on Agents and Data Mining Interaction, ADMI 2011, Taipei, Taiwan, May 2– 6, 2011, Revised Selected Papers, volume 7103 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27608-3 (print), 3-642-27609-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27609-5.

Cohen-Boulakia:2007:DIL

[CBT07] Sarah Cohen-Boulakia and Val Tannen, editors. Data Integration in the Life Sciences: 4th International Workshop, DILS 2007, Philadelphia, PA, USA, June 27–29, 2007. Proceedings, volume 4544 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73254-3 (print), 3-540-73255-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73255-6.

Cornuejols:1999:IPC

[CBW99] Gérard Cornuéjols, Rainer E.

Burkard, and Gerhard J. Woeginger, editors. Integer programming and combinatorial optimization: 7th International IPCO Conference, Graz, Austria, June 9–11, 1999: Proceedings, volume 1610 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66019-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T57.74 .I545 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1610.htm;
http://www.springerlink.
com/content/978-3-540-
66019-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1610.

Charpin:1991:EIS

[CC91] P. Charpin and G. Cohen, editors. Eurocode ’90: International Symposium on coding theory and applications, Udine, Italy, November 5–9, 1990: proceedings, volume 514 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54303-1, 0-387-54303-

1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .E95 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0514.htm;
 http://www.springerlink.
com/content/978-3-540-
54303-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=514.

Calmet:1993:AIS

[CC93] Jacques Calmet and J. A. Campbell, editors. Artificial intelligence and symbolic mathematical computing: International Conference AISMC-1, Karlsruhe, Germany, August 1992: proceedings, volume 737 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-57322-4 (U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I5488 1992. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0737.
htm;   http://www.springerlink.
com/content/978-0-387-
57322-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=737.

Calmet:1995:ISM

[CC95] Jacques Calmet and J. A. Campbell, editors. Integrating symbolic mathematical computation and artificial intelligence: Second International Conference, AISMC-2, Cambridge, United Kingdom, August 3–5, 1994: Selected Papers, volume 958 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60156-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 I567 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0958.
htm;   http://www.springerlink.
com/content/978-3-540-
60156-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=958.

Chlebus:1997:FCT

[CC97a] Bogdan Chlebus and Ludwik Czaja, editors. Fundamentals of computation theory: 11th International Symposium, FCT ’97, Krakow, Poland, September 1– 3, 1997: proceedings, volume 1279 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63386-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5.F367 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1279.
htm;   http://www.springerlink.
com/content/978-3-540-
63386-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1279.

Costa:1997:PAI

[CC97b] Ernesto Costa and Amilcar Cardoso, editors. Progress in artificial intelligence: 8th Portuguese Conference on Artificial Intelligence, EPIA ’97, Coimbra, Portugal, October 6–9, 1997: proceedings, volume 1323 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63586-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P67 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1323.htm;
 http://www.springerlink.
com/content/978-3-540-
63586-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1323.

Cachin:2004:ACE

[CC04] Christian Cachin and Jan Camenisch, editors. Advances in Cryptology—EUROCRYPT 2004: International Conference on the Theory and Applications of Cryptographic Techniques, Interlaken, Switzerland, May 2–6, 2004: Proceedings, volume 3027 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21935-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3027.htm;
 http://www.springerlink.
com/content/978-3-540-
21935-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=3027;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97182.

Candan:2005:AMI

[CC05] K. Selçuk Candan and Augusto Celentano, editors. Advances in multimedia information systems: 11th international workshop, MIS 2005, Sorrento, Italy, September 19–21, 2005, proceedings, volume 3665 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28792-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28792-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3665.

Cerqueira:2007:MAI

[CC07] Renato Cerqueira and Roy H. Campbell, editors. Middleware 2007: ACM/IFIP/USENIX 8th International Middleware Conference, Newport Beach, CA, USA, November 26–30, 2007. Proceedings, volume 4834 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76777-0 (print), 3-540-76778-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76778-7.

Chen:2009:TAM

[CC09a] Jianer Chen and S. Barry Cooper, editors. Theory and Applications of Models of Computation: 6th Annual Conference, TAMC 2009, Changsha, China, May 18–22, 2009. Proceedings, volume 5532 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02016-X (print), 3-642-02017-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02017-9.

Cotta:2009:ECC

[CC09b] Carlos Cotta and Peter Cowling, editors. Evolutionary Computation in Combinatorial Optimization: 9th European Conference, Evo-COP 2009, Tübingen, Germany, April 15– 17, 2009. Proceedings, volume 5482 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01008-3 (print), 3-642-01009-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01009-5.

Chen:2012:ODL

[CC12] Hsin-Hsi Chen and Gobinda Chowdhury, editors. The Outreach of Digital Libraries: A Globalized Resource Network: 14th International Conference on Asia-Pacific Digital Libraries, ICADL 2012, Taipei, Taiwan, November 12–15, 2012, Proceedings, volume 7634 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34751-7 (print), 3-642-34752-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34752-8.

Choy:2012:RCA

[CCAS12] Oliver C. S. Choy, Ray C. C. Cheung, Peter Athanas, and Kentaro Sano, editors. Reconfigurable Computing: Architectures, Tools and Applications: 8th International Symposium, ARC 2012, Hong Kong, China, March 19–23, 2012. Proceedings, volume 7199 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28364-0 (print), 3-642-28365-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28365-9.

Chrisment:2011:MDN

[CCBW11] Isabelle Chrisment, Alva Couch, Rémi Badonnel, and Martin Waldburger, editors. Managing the Dynamics of Networks and Services: 5th International Conference on Autonomous Infrastructure, Management, and Security, AIMS 2011, Nancy, France, June 13– 17, 2011. Proceedings, volume 6734 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21483-5 (print), 3-642-21484-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21484-4.

Calude:2004:DLT

[CCD04] Cristian S. Calude, Elena Calude, and Michael J. Dinneen, editors. Developments in language theory: 8th International Conference, DLT 2004, Auckland, New Zealand, December 13–17, 2004: Proceedings, volume 3340 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24014-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 A1 L43 3340 (LC). URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24014-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3340;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b103739.

Cham:2006:AMMa

[CCD^{+}06a] Tat-Jen Cham, Jianfei Cai, Chitra Dorai, Deepu Rajan, Tat-Seng Chua, and Liang-Tien Chia, editors. Advances in Multimedia Modeling: 13th International Multimedia Modeling Conference, MMM 2007, Singapore, January 9–12, 2007. Proceedings, Part I, volume 4351 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69421-8 (print), 3-540-69423-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69423-6.

Cham:2006:AMMb

[CCD^{+}06b] Tat-Jen Cham, Jianfei Cai, Chitra Dorai, Deepu Rajan, Tat-Seng Chua, and Liang-Tien Chia, editors. Advances in Multimedia Modeling: 13th International Multimedia Modeling Conference, MMM 2007, Singapore, January 9–12, 2007. Proceedings, Part II, volume 4352 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69428-5 (print), 3-540-69429-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69429-8.

Calude:2009:UCI

[CCD^{+}09] Cristian S. Calude, José Félix Costa, Nachum Dershowitz, Elisabete Freire, and Grzegorz Rozenberg, editors. Unconventional Computation: 8th International Conference, UC 2009, Ponta Delgada, Portugal, September 7–11, 2009. Proceedings, volume 5715 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03744-5 (print), 3-642-03745-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03745-0.

Coello:2012:PPSa

[CCD^{+}12a] Carlos A. Coello Coello, Vincenzo Cutello, Kalyanmoy Deb, Stephanie Forrest, Giuseppe Nicosia, and Mario Pavone, editors. Parallel Problem Solving from Nature — PPSN XII: 12th International Conference, Taormina, Italy, September 1–5, 2012, Proceedings, Part I, volume 7491 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32936-5 (print), 3-642-32937-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32937-1.

Coello:2012:PPSb

[CCD^{+}12b] Carlos A. Coello Coello, Vincenzo Cutello, Kalyanmoy Deb, Stephanie Forrest, Giuseppe Nicosia, and Mario Pavone, editors. Parallel Problem Solving from Nature — PPSN XII: 12th International Conference, Taormina, Italy, September 1–5, 2012, Proceedings, Part II, volume 7492 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32963-2 (print), 3-642-32964-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32964-7.

Cheung:2003:FPL

[CCdS03] Peter Y. K. Cheung, George A. Constantinides, and José T. de Sousa, editors. Field-Programmable Logic and Applications: 13th International Conference, FPL 2003, Lisbon, Portugal, September 1–3, 2003: Proceedings, volume 2778 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40822-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.G36 I48 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2778.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40822-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2778;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12007.

Calude:2008:UCI

[CCF^{+}08] Cristian S. Calude, José Félix Costa, Rudolf Freund, Marion Oswald, and Grzegorz Rozenberg, editors. Unconventional Computing: 7th International Conference, UC 2008 Vienna, Austria, August 25–28, 2008. Proceedings, volume 5204 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85193-3 (print), 3-540-85194-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85194-3.

Christophides:2008:SWO

[CCG08] Vassilis Christophides, Martine Collard, and Claudio Gutierrez, editors. Semantic Web, Ontologies and Databases: VLDB Workshop, SWDB-ODBIS 2007, Vienna, Austria, September 24, 2007, Revised Selected Papers, volume 5005 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70959-2 (print), 3-540-70960-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70960-2.

Chen:2002:AMI

[CCH02] Yung-Chang Chen, Longwen Chang, and Chiou-Ting Hsu, editors. Advances in multimedia information processing — PCM 2002: third IEEE Pacific Rim Conference on Multimedia, Hsinchu, Taiwan, December 16– 18, 2002: Proceedings, volume 2532 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00262-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I333 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2532.htm;
 http://link.springer.
de/link/service/series/
0558/tocs/t2532.htm;
 http://www.springerlink.
com/content/978-3-540-
00262-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2532.

Also available via the World Wide Web.

Chen:2010:ISI

[CChL^{+}10] Hsinchun Chen, Michael Chau, Shu hsing Li, Shalini Urs, Srinath Srinivasa, and G. Alan Wang, editors. Intelligence and Security Informatics: Pacific Asia Workshop, PAISI 2010, Hyderabad, India, June 21, 2010. Proceedings, volume 6122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13600-1 (print), 3-642-13601-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13601-6.

Cesar:2007:ITS

[CCJ07] Pablo Cesar, Konstantinos Chorianopoulos, and Jens F. Jensen, editors. Interactive TV: a Shared Experience: 5th European Conference, EuroITV 2007, Amsterdam, The Netherlands, May 24–25, 2007. Proceedings, volume 4471 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72558-X (print), 3-540-72559-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72559-6.

Camenisch:2007:IHI

[CCJS07] Jan L. Camenisch, Christian S. Collberg, Neil F. Johnson, and Phil Sallee, editors. Information Hiding: 8th International Workshop, IH 2006, Alexandria, VA, USA, July 10–12, 2006. Revised Selected Papers, volume 4437 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74123-2 (print), 3-540-74124-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74124-4.

Clarke:1995:BTS

[CCK95] Anne Clarke, Mario Campolargo, and Nikos Karatzas, editors. Bringing telecommunication services to the people — IS and N ’95: Third International Conference on Intelligence in Broadband Services and Networks, Heraklion, Crete, Greece, October 16–19, 1995: proceedings, volume 998 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60479-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.A1 I58 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0998.htm;
http://www.springerlink.
com/content/978-3-540-
60479-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=998.

Cha:2008:ATV

[CCK^{+}08] Sungdeok (Steve) Cha, Jin-Young Choi, Moonzoo Kim, Insup Lee, and Mahesh Viswanathan, editors. Automated Technology for Verification and Analysis: 6th International Symposium, ATVA 2008, Seoul, Korea, October 20–23, 2008. Proceedings, volume 5311 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88386-X (print), 3-540-88387-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88387-6.

Chang:2002:RAV

[CCL02] S. K. Chang, Zen Chen, and Suh-Yin Lee, editors. Recent advances in visual information systems: 5th International Conference, VISUAL 2002, Hsin Chu, Taiwan, March 11–13, 2002: Proceedings, volume 2314 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43358-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.I6 V56 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2314.
htm;   http://www.springerlink.
com/content/978-3-540-
43358-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2314.

Cai:2006:TAM

[CCL06] Jin-Yi Cai, S. Barry Cooper, and Angsheng Li, editors. Theory and Applications of Models of Computation: Third International Conference, TAMC 2006, Beijing, China, May 15–20, 2006, Proceedings, volume 3959 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34021-1 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34021-8.

Christianson:1998:SPI

[CCLR98] Bruce Christianson, Bruno Crispo, T. Mark A. Lomas, and Michael Roe, editors. Security protocols: 5th International Workshop, Paris, France, April 7–9, 1997: proceedings, volume 1361 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64040-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25S44 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1361.htm;
http://www.springerlink.
com/content/978-3-540-
64040-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1361.

Chen:2004:DLI

[CCM^{+}04] Zhaoneng Chen, Hsinchun Chen, Qihao Miao, Yuxi Fu, Edward Fox, and Ee peng Lim, editors. Digital Libraries: International Collaboration and Cross-Fertilization: 7th International Conference on Asian Digital Libraries, ICADL 2004, Shanghai, China, December 13–17, 2004: Proceedings, volume 3334 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24030-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4080 .I578 2004. URL http:

//www.springerlink.
com/content/978-3-540-
24030-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3334;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104284.

Clark:2008:GIA

[CCM08] Alexander Clark, François Coste, and Laurent Miclet, editors. Grammatical Inference: Algorithms and Applications: 9th International Colloquium, ICGI 2008 Saint-Malo, France, September 22–24, 2008 Proceedings, volume 5278 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88008-9 (print), 3-540-88009-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
88009-7.

Christianson:2002:SPI

[CCMR02] Bruce Christianson, Bruno Crispo, James A. Malcolm, and Michael Roe, editors. Security Protocols: 9th International Workshop Cambridge, UK, April 25– 27, 2001. Revised Papers, volume 2467 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44263-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 S443 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2467.
htm;   http://www.springerlink.
com/content/978-3-540-
44263-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2467.

Christianson:2003:SPI

[CCMR03] Bruce Christianson, Bruno Crispo, James A. Malcolm, and Michael Roe, editors. Security Protocols: 10th International Workshop, Cambridge, UK, April 17– 19, 2002: Revised Papers, volume 2845 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20830-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I574 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2845.
htm;   http://www.springerlink.
com/content/978-3-540-
20830-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2845;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94819.

Christianson:2005:SPI

[CCMR05] Bruce Christianson, Bruno Crispo, James A. Malcolm, and Michael Roe, editors. Security Protocols: 11th International Workshop, Cambridge, UK, April 2–4, 2003, Revised Selected Papers, volume 3364 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28389-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28389-8;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3364.

Christianson:2006:SPI

[CCMR06] Bruce Christianson, Bruno Crispo, James A. Malcolm, and Michael Roe, editors. Security Protocols: 12th International Workshop, Cambridge, UK, April 26–28, 2004. Revised Selected Papers, volume 3957 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-40925-4 (print), 3-540-40926-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
40926-7.

Christianson:2007:SPI

[CCMR07] Bruce Christianson, Bruno Crispo, James A. Malcolm, and Michael Roe, editors. Security Protocols: 13th International Workshop, Cambridge, UK, April 20–22, 2005, Revised Selected Papers, volume 4631 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77155-7 (print), 3-540-77156-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77156-2.

Christianson:2009:SPI

[CCMR09] Bruce Christianson, Bruno Crispo, James A. Malcolm, and Michael Roe, editors. Security Protocols: 14th International Workshop, Cambridge, UK, March 27–29, 2006, Revised Selected Papers, volume 5087 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04903-6 (print), 3-642-04904-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04904-0.

Christianson:2010:SPI

[CCMR10] Bruce Christianson, Bruno Crispo, James A. Malcolm, and Michael Roe, editors. Security Protocols: 15th International Workshop, Brno, Czech Republic, April 18–20, 2007. Revised Selected Papers, volume 5964 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17772-7 (print), 3-642-17773-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17773-6.

Christianson:2011:SPXb

[CCMS11] Bruce Christianson, Bruno Crispo, James Malcolm, and Frank Stajano, editors. Security Protocols XIX: 19th International Workshop, Cambridge, UK, March 28–30, 2011, Revised Selected Papers, volume 7114 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25866-2 (print), 3-642-25867-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25867-1.

Chignell:2010:SIC

[CCNY10] Mark Chignell, James Cordy, Joanna Ng, and Yelena Yesha, editors. The Smart Internet: Current Research and Future Applications, volume 6400 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16598-2 (print), 3-642-16599-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16599-3.

Calmet:1996:AIS

[CCP96] Jacques Calmet, J. A. Campbell, and Jochen Pfalzgraf, editors. Artificial intelligence and symbolic mathematical computation: international conference, AISMC-3, Steyr, Austria, September 23–25, 1996, proceedings, volume 1138 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61732-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A766 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1138.htm;
 http://www.springerlink.
com/content/978-3-540-
61732-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1138.

Third International Conference on Artificial Intelligence and Symbolic Mathematical Computation (AISMC-3), held in Steyr (Austria), September 23–25, 1996.

Cheong:2010:ACIa

[CCP10a] Otfried Cheong, Kyung-Yong Chwa, and Kunsoo Park, editors. Algorithms and Computation: 21st International Symposium, ISAAC 2010, Jeju Island, Korea, December 15–17, 2010, Proceedings, Part I, volume 6506 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17516-3 (print), 3-642-17517-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17517-6.

Cheong:2010:ACIb

[CCP10b] Otfried Cheong, Kyung-Yong Chwa, and Kunsoo Park, editors. Algorithms and Computation: 21st International Symposium, ISAAC 2010, Jeju, Korea, December 15–17, 2010, Proceedings, Part II, volume 6507 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17513-9 (print), 3-642-17514-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17514-5.

Cerri:2012:ITS

[CCPP12] Stefano A. Cerri, William J. Clancey, Giorgos Papadourakis, and Kitty Panourgia, editors. Intelligent Tutoring Systems: 11th International Conference, ITS 2012, Chania, Crete, Greece, June 14–18, 2012. Proceedings, volume 7315 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30949-6 (print), 3-642-30950-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30950-2.

Cousineau:1986:CFP

[CCR86] Guy Cousineau, P.-L. Curien, and B. Robinet, editors. Combinators and functional programming languages: Thirteenth Spring School of the LITP, Val d’Ajol, France, May 6–10, 1985, proceedings, volume 242 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-17184-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .L331 1985. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0242.htm;
 http://www.springerlink.
com/content/978-0-387-
17184-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=242.

Callan:2004:DMI

[CCS04] Jamie Callan, Fabio Crestani, and Mark Sanderson, editors. Distributed Multimedia Information Retrieval: SI-GIR 2003 Workshop on Distributed Information Retrieval, Toronto, Canada, August 1, 2003: Revised, Selected, and Invited Papers, volume 2924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20875-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D584 2004.

URL http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t2924.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20875-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2924;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95004.

Cortesi:2012:CIS

[CCSW12] Agostino Cortesi, Nabendu Chaki, Khalid Saeed, and Sławomir Wierzchoń, editors. Computer Information Systems and Industrial Management: 11th IFIP TC 8 International Conference, CISIM 2012, Venice, Italy, September 26– 28, 2012. Proceedings, volume 7564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33259-5 (print), 3-642-33260-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33260-9.

Cerri:2000:AIM

[CD00a] Stefano A. Cerri and Danail Dochev, editors. Artificial Intelligence:

Methodology, Systems, and Applications: 9th International Conference, AIMSA 2000, Varna, Bulgaria, September 20– 23, 2000. Proceedings, volume 1904 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41044-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 I554 2000; QA267.A1 L43 no.1904.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1904.
htm;   http://www.springerlink.
com/content/978-3-540-
41044-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1904.

Chhabra:2000:GRR

[CD00b] Atul K. Chhabra and Dov Dori, editors. Graphics Recognition. Recent Advances: Third International Workshop, GREC’99, Jaipur, India, September 26–27, 1999. Selected Papers, volume 1941 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41222-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1941. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1941.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41222-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1941.

Cussens:2000:LLL

[CD00c] James Cussens and Sašo Džeroski. Learning Language in Logic, volume 1925 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41145-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 299 pp. LCCN QA76.9.L63 L42 2000.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1925.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        41145-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1925.

Cotronis:2001:RAP

[CD01] Yiannis Cotronis and J. J. Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 8th European PVM/MPI Users’ Group Meeting, Santorini/Thera, Greece, September 23–26, 2001: Proceedings, volume 2131 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42609-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E975 2001; QA267.A1 L43 no.2131. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2131.htm;
 http://www.springerlink.
com/content/978-3-540-
42609-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2131.

Ciardo:2005:ATP

[CD05] Gianfranco Ciardo and Philippe Darondeau, editors. Applications and theory of Petri nets 2005: 26th international conference, ICAPTN 2005, Miami, USA, June 20– 25, 2005, proceedings, volume 3536 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26301-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26301-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3536;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136988.

Cavazza:2007:VSU

[CD07] Marc Cavazza and Stéphane Donikian, editors. Virtual Storytelling. Using Virtual Reality Technologies for Storytelling: 4th International Conference, ICVS 2007, Saint-Malo, France, December 5–7, 2007. Proceedings, volume 4871 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77037-2 (print), 3-540-77039-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77039-8.

Cavalcanti:2009:FFM

[CD09] Ana Cavalcanti and Dennis R. Dams, editors. FM 2009: Formal Methods: Second World Congress, Eindhoven, The Netherlands, November 2– 6, 2009. Proceedings, volume 5850 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05088-3 (print), 3-642-05089-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
05089-3.

Calamoneri:2010:ACI

[CD10] Tiziana Calamoneri and Josep Diaz, editors. Algorithms and Complexity: 7th International Conference, CIAC 2010, Rome, Italy, May 26– 28, 2010. Proceedings, volume 6078 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13072-0 (print), 3-642-13073-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        13073-1.

Cuzzocrea:2011:DWK

[CD11] Alfredo Cuzzocrea and Umeshwar Dayal, editors. Data Warehousing and Knowledge Discovery: 13th International Conference, DaWaK 2011, Toulouse, France, August 29– September 2,2011. Proceedings, volume 6862 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23543-3 (print), 3-642-23544-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23544-3.

Cuzzocrea:2012:DWK

[CD12] Alfredo Cuzzocrea and Umeshwar Dayal, editors. Data Warehousing and Knowledge Discovery: 14th International Conference, DaWaK 2012, Vienna, Austria, September 3– 6, 2012. Proceedings, volume 7448 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32583-1 (print), 3-642-32584-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32584-7.

Cruz:2006:SWI

[CDA^{+}06] Isabel Cruz, Stefan Decker, Dean Allemang, Chris Preist, Daniel Schwabe, Peter Mika, Mike Uschold, and Lora M. Aroyo, editors. The Semantic Web — ISWC 2006: 5th International Semantic Web Conference, ISWC 2006, Athens, GA, USA, November 5–9, 2006. Proceedings, volume 4273 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49029-9 (print), 3-540-49055-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49055-5.

Christensen-Dalsgaard:2008:RAT

[CDCJL08] Birte Christensen-Dalsgaard, Donatella Castelli, Bolette Ammitzbøll Jurik, and Joan Lippincott, editors. Research and Advanced Technology for Digital Libraries: 12th European Conference, ECDL 2008, Aarhus, Denmark, September 14– 19, 2008. Proceedings, volume 5173 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87598-0 (print), 3-540-87599-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87599-4.

Carette:2009:ICM

[CDCW09] Jacques Carette, Lucas Dixon, Claudio Sacerdoti Coen, and Stephen M. Watt, editors. Intelligent Computer Mathematics: 16th Symposium, Calculemus 2009, 8th International Conference, MKM 2009, Held as Part of CICM 2009, Grand Bend, Canada, July 6–12, 2009. Proceedings, volume 5625 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02613-3 (print), 3-642-02614-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02614-0.

Clavel:2007:AAM

[CDE^{+}07] Manuel Clavel, Francisco Durán, Steven Eker, Patrick Lincoln, Narciso Martí-Oliet, José Meseguer, and Carolyn Talcott, editors. All About Maude — A High-Performance Logical Framework: How to Specify, Program and Verify Systems in Rewriting Logic, volume 4350 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71940-7 (print), 3-540-71999-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 665 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71999-1.

Costabile:2011:EUD

[CDFP11] Maria Francesca Costabile, Yvonne Dittrich, Gerhard Fischer, and Antonio Piccinno, editors. End-User Development: Third International Symposium, IS-EUD 2011, Torre Canne (BR), Italy, June 7– 10, 2011. Proceedings, volume 6654 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21529-7 (print), 3-642-21530-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21530-8.

Csirik:1989:FCT

[CDG89] J. Csirik, Janos Demetrovics, and F. Gecseg, editors. Fundamentals of computation theory: International Conference, FCT’89, Szeged, Hungary, August 21– 25, 1989: proceedings, volume 380 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51498-8 (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 I5731 1989. DM69.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0380.htm;
 http://www.springerlink.
com/content/978-0-387-
51498-7;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=380.

Cavalcanti:2010:TAC

[CDGW10] Ana Cavalcanti, David Deharbe, Marie-Claude Gaudel, and Jim Woodcock, editors. Theoretical Aspects of Computing — ICTAC 2010: 7th International Colloquium, Natal, Rio Grande do Norte, Brazil, September 1–3, 2010. Proceedings, volume 6255 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14807-7 (print), 3-642-14808-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14808-8.

Cooper:2012:HWC

[CDL12] S. Barry Cooper, Anuj Dawar, and Benedikt Löwe, editors. How the World Computes: Turing Centenary Conference and 8th Conference on Computability in Europe, CiE 2012, Cambridge, UK, June 18–23, 2012. Proceedings, volume 7318 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-30869-4. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30869-7.

Cheng:2009:SES

[CdLG^{+}09] Betty H. C. Cheng, Rogério de Lemos, Holger Giese, Paola Inverardi, and Jeff Magee, editors. Software Engineering for Self-Adaptive Systems, volume 5525 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02160-3 (print), 3-642-02161-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02161-9.

Casimiro:2010:ADS

[CdLG10] Antonio Casimiro, Rogério de Lemos, and Cristina Gacek, editors. Architecting Dependable Systems VII, volume 6420 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17244-X (print), 3-642-17245-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17245-8.

Chen:2006:ISP

[CDLZ06] Kefei Chen, Robert Deng, Xuejia Lai, and Jianying Zhou, editors. Information Security Practice and Experience: Second International Conference, IS-PEC 2006, Hangzhou, China, April 11–14, 2006. Proceedings, volume 3903 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33052-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33052-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3903.

Crestani:2004:MUI

[CDM04] Fabio Crestani, Mark Dunlop, and Stefano Mizzaro, editors. Mobile and Ubiquitous Information Access: Mobile HCI 2003 International Workshop, Udine, Italy, September 8, 2003: Revised and Invited Papers, volume 2954 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21003-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK6570.M6 W67 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2954.
htm;   http://www.springerlink.
com/content/978-3-540-
21003-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2954;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95483.

Collier:2012:PMA

[CDN12] Rem Collier, Jürgen Dix, and Peter Novák, editors. Programming Multi-Agent Systems: 8th International Workshop, ProMAS 2010, Toronto, ON, Canada, May 11, 2010. Revised Selected Papers, volume 6599 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28938-X (print), 3-642-28939-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
28939-2.

Cotronis:2011:RAM

[CDND11] Yiannis Cotronis, Anthony Danalis, Dimitrios S. Nikolopoulos, and Jack Dongarra, editors. Recent Advances in the Message Passing Interface: 18th European MPI Users’ Group Meeting, EuroMPI 2011, Santorini, Greece, September 18–21, 2011. Proceedings, volume 6960 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24448-3 (print), 3-642-24449-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24449-0.

Coquand:2000:TPP

[CDNS00] Thierry Coquand, Peter Dybjer, Bengt Nordström, and Jan Smith, editors. Types for Proofs and Programs: International Workshop, TYPES’99, Lökeberg, Sweden, June 12–16, 1999. Selected Papers, volume 1956 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41517-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T96 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1956.htm;
 http://www.springerlink.
com/content/978-3-540-
41517-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1956.

Calude:2002:UMC

[CDP02] Cristian S. Calude, M. J. Dinneen, and Ferdinand Peper, editors. Unconventional Models of Computation: Third International Conference, UMC 2002, Kobe, Japan, October 15–19, 2002. Proceedings, volume 2509 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44311-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S63 I574 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2509.
htm;   http://www.springerlink.
com/content/978-3-540-
44311-7;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2509.

Calude:2005:UCI

[CDP^{+}05] Cristian S. Calude, Michael J.

Dinneen, Gheorghe Paun, Mario J. Pérez-Jiménez, and Grzegorz Rozenberg, editors. Unconventional computation: 4th international conference, UC 2005, Sevilla, Spain, October 3–7, 2005: proceedings, volume 3699 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29100-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S63 I574 2005.

URL http://www.springerlink.
com/content/978-3-540-
29100-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3699.

Calude:2006:UCI

[CDP^{+}06] Cristian S. Calude, Michael J.

Dinneen, Gheorghe Păun, Grzegorz Rozenberg, and Susan Stepney, editors. Unconventional Computation: 5th International Conference, UC 2006, York, UK, September 4–8, 2006. Proceedings, volume 4135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38593-2 (print), 3-540-38594-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        38594-3.

Chaudhuri:2006:DCN

[CDPT06] Soma Chaudhuri, Samir R. Das, Himadri S. Paul, and Srikanta Tirthapura, editors. Distributed Computing and Networking: 8th International Conference, ICDCN 2006, Guwahati, India, December 27–30, 2006. Proceedings, volume 4308 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68139-6 (print), 3-540-68140-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68140-3.

Cress:2009:LSM

[CDS09] Ulrike Cress, Vania Dimitrova, and Marcus Specht, editors. Learning in the Synergy of Multiple Disciplines: 4th European Conference on Technology Enhanced Learning, EC-TEL 2009 Nice, France, September 29–October 2, 2009 Proceedings, volume 5794 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04635-5 (print), 3-642-04636-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04636-0.

Concas:2007:APS

[CDSS07] Giulio Concas, Ernesto Damiani, Marco Scotto, and Giancarlo Succi, editors. Agile Processes in Software Engineering and Extreme Programming: 8th International Conference, XP 2007, Como, Italy, June 18–22, 2007. Proceedings, volume 4536 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73100-8 (print), 3-540-73101-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73101-6.

Crowley:2011:CVS

[CDT11] James L. Crowley, Bruce A. Draper, and Monique Thonnat, editors. Computer Vision Systems: 8th International Conference, ICVS 2011, Sophia Antipolis, France, September 20–22, 2011. Proceedings, volume 6962 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23967-6 (print), 3-642-23968-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23968-7.

Calude:2003:DMT

[CDV03] Cristian S. Calude, Michael J. Dinneen, and Vincent Vajnovszki, editors. Discrete Mathematics and Theoretical Computer Science: 4th International Conference, DMTCS 2003, Dijon, France, July 7–12, 2003, Proceedings, volume 2731 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40505-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA164 .D62 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2731.htm;
 http://www.springerlink.
com/content/978-3-540-
40505-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2731.

Conradi:1986:APE

[CDW86] Reidar Conradi, Tor M.

Didriksen, and Dag H. Wanvik, editors. Advanced programming environments: proceedings of an International Workshop, Trondheim, Norway, June 16–18, 1986, volume 244 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-17189-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .A33271 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0244.htm;
 http://www.springerlink.
com/content/978-0-387-
17189-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=244.

Papers presented at the International Workshop on Advanced Programming Environments, organized by IFIP Working Group 2.4 on Systems Programming Languages in cooperation with ACM SIGPLAN/SIGSOFT, sponsored by Division of Computer Science (DCS), NTH and other organizations.

Christiansen:2011:FQA

[CDY^{+}11] Henning Christiansen, Guy De Tré, Adnan Yazici, Slawomir Zadrozny, Troels Andreasen, and Henrik Legind Larsen, editors. Flexible Query Answering Systems: 9th International Conference, FQAS 2011, Ghent, Belgium, October 26– 28, 2011 Proceedings, volume 7022 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24763-6 (print), 3-642-24764-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24764-4.

Camurati:1995:CHD

[CE95] Paolo Camurati and Hans Eveking, editors. Correct hardware design and verification methods: IFIP WG 10.5 advanced research working conference, CHARME 95, Frankfurt/Main, Germany, October 2–4, 1995: proceedings, volume 987 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60385-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.75 .A39 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0987.htm;
 http://www.springerlink.
com/content/978-3-540-
60385-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=987.

Czarnecki:2000:GCB

[CE00] Krzysztof Czarnecki and Ulrich Eisenecker, editors. Generative and component-based software engineering: First International Symposium, GCSE’99, Erfurt, Germany, September 28– 30, 1999: Revised Papers, volume 1799 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41172-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.758 .G418 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1799.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41172-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1799.

Caplinskas:2001:ADI

[CE01] Albertas Caplinskas and J. Eder, editors. Advances in databases and information systems: 5th East European Conference, ADBIS 2001, Vilnius, Lithuania, September 25–28, 2001: Proceedings, volume 2151 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42555-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 A3396 2001; QA267.A1 L43 no.2151. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2151.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42555-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2151.

Cea:1976:OTMa

[Cea76a] Jean Cea, editor. Optimization techniques: modeling and optimization in the service of man: proceedings, 7th IFIP conference, Nice, September 8–12, 1975. Part I, volume 40 of Series, I.F.I.P. TC 7 optimization conferences and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1976. CO-DEN LNCSD9. ISBN 0-387-07622-0 (New York), 3-540-07622-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA402.5 .I1731 1975. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0040.htm;
http://www.springerlink.
com/content/978-0-387-
07622-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=40.

Sponsored by the IFIP Technical Committee on Optimization (TC 7).

Cea:1976:OTMb

[Cea76b] Jean Cea, editor. Optimization techniques: modeling and optimization in the service of man: proceedings, 7th IFIP conference, Nice, September 8–12, 1975.

Part II, volume 41 of Series, I.F.I.P. TC 7 optimization conferences and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1976. CO-DEN LNCSD9. ISBN 0-387-07623-9 (New York), 3-540-07623-9 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA402.5 .I1731 1975. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0041.htm;
 http://www.springerlink.
com/content/978-0-387-
07623-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=41.

Sponsored by the IFIP Technical Committee on Optimization (TC 7).

Cebulla:2008:OOT

[Ceb08] Michael Cebulla, editor. Object-Oriented Technology. ECOOP 2007 Workshop Reader: ECOOP 2007 Workshops, Berlin, Germany, July 30–31, 2007, Final Reports, volume 4906 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78194-3 (print), 3-540-78195-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78195-0.

Cuny:1996:GGT

[CEER96] J. (Janice) Cuny, Hartmut Ehrig, Gregor Engels, and Grzegorz Rozenberg, editors. Graph Grammars and Their Application to Computer Science: 5th International Workshop, Williamsburg, VA, USA, November 13–18, 1994: Selected Papers, volume 1073 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61228-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .G68 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1073.htm;
http://www.springerlink.
com/content/978-3-540-
61228-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1073.

Chollet:2005:NSM

[CEFZM05] Gérard Chollet, Anna Esposito, Marcos Faundez-Zanuy, and Maria Marinaro, editors. Nonlinear Speech Modeling and Applications: Advanced Lectures and Revised Selected Papers, volume 3445 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27441-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.S65. URL http:

//www.springerlink.
com/content/978-3-540-
27441-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3445.

Chen:2011:INE

[CEK11] Ning Chen, Edith Elkind, and Elias Koutsoupias, editors. Internet and Network Economics: 7th International Workshop, WINE 2011, Singapore, December 11–14, 2011. Proceedings, volume 7090 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25509-4 (print), 3-642-25510-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25510-6.

Corradini:2002:GTF

[CEKR02] Andrea Corradini, Hartmut Ehrig, Hans-Jörg Kreowski, and Grzegorz Rozenberg, editors. Graph Transformation: First International Conference, ICGT 2002, Barcelona, Spain, October 7–12, 2002. Proceedings, volume 2505 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44310-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .I34 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2505.htm;
http://www.springerlink.
com/content/978-3-540-
44310-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2505.

Corradini:2006:GTT

[CEM^{+}06] Andrea Corradini, Hartmut Ehrig, Ugo Montanari, Leila Ribeiro, and Grzegorz Rozenberg, editors. Graph Transformations: Third International Conference, ICGT 2006 Natal, Rio Grande do Norte, Brazil, September 17–23, 2006 Proceedings, volume 4178 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38870-2 (print), 3-540-38872-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38872-2.

Cheng:2011:WAS

[CEQ^{+}11] Yu Cheng, Do Young Eun, Zhiguang Qin, Min Song, and Kai Xing, editors. Wireless Algorithms, Systems, and Applications: 6th International Conference, WASA 2011, Chengdu, China, August 11–13, 2011. Proceedings, volume 6843 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23489-5 (print), 3-642-23490-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23490-3.

Claus:1979:GGT

[CER79] Volker Claus, Hartmut Ehrig, and Grzegorz Rozenberg, editors. Graph-grammars and their application to computer science and biology: International Workshop, Bad Honnef, October 30–November 3, 1978, volume 73 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CODEN LNCSD9. ISBN 0-387-09525-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .G71. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0073.htm;
 http://www.springerlink.
com/content/978-0-387-
09525-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=73.

Cerioli:2005:FAS

[Cer05] Maura Cerioli, editor.

Fundamental Approaches to Software Engineering: 8th International Conference, FASE 2005, held as part of the Joint Conferences on Theory and Practice of Software, ETAPS 2005, Edinburgh, UK, April 4–8, 2005: Proceedings, volume 3442 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25420-X (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F44 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25420-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3442;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b107062.

Cervesato:2007:ACS

[Cer07] Iliano Cervesato, editor. Advances in Computer Science — ASIAN 2007. Computer and Network Security: 12th Asian Computing Science Conference, Doha, Qatar, December 9–11, 2007. Proceedings, volume 4846 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76927-7 (print), 3-540-76929-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76929-3.

Chang:1980:PIS

[CF80] S. K. Chang and K. S. Fu. Pictorial information systems, volume 80 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-09757-0. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 445 pp. LCCN TA1632 .P5. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0080.htm;
 http://www.springerlink.
com/content/978-0-387-
09757-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=80.

Cortesi:1999:SAI

[CF99] Agostino Cortesi and Gilberto Filé, editors. Static analysis: 6th International Symposium, SAS’99, Venice, Italy, September 22–24, 1999: Proceedings, volume 1694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66459-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1694.htm;
 http://www.springerlink.
com/content/978-3-540-
66459-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1694.

Carter:2000:LCP

[CF00a] Larry Carter and Jeanne Ferrante, editors. Languages and compilers for parallel computing: 12th International Workshop, LCPC’99, La Jolla, CA, USA, August 4–6, 1999: Proceedings, volume 1863 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67858-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .L36 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1863.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67858-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1863.

Cussens:2000:ILP

[CF00b] James Cussens and Alan Frisch, editors. Inductive logic programming: 10th International Conference, ILP 2000, London, UK, July 2000: Proceedings, volume 1866 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67795-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1866. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1866.htm;
http://www.springerlink.
com/content/978-3-540-
67795-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1866.

Chen:2009:PEC

[CF09a] Jianer Chen and Fedor V. Fomin, editors. Parameterized and Exact Computation: 4th International Workshop, IWPEC 2009, Copenhagen, Denmark, September 10–11, 2009, Revised Selected Papers, volume 5917 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-11268-4 (print), 3-642-11269-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11269-0.

Cofer:2009:FMI

[CF09b] Darren Cofer and Alessandro Fantechi, editors. Formal Methods for Industrial Critical Systems: 13th International Workshop, FMICS 2008, L’Aquila, Italy, September 15–16, 2008, Revised Selected Papers, volume 5596 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03239-7 (print), 3-642-03240-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03240-0.

Coyle:2010:AIC

[CF10] Lorcan Coyle and Jill Freyne, editors. Artificial Intelligence and Cognitive Science: 20th Irish Conference, AICS 2009, Dublin, Ireland, August 19–21, 2009, Revised Selected Papers, volume 6206 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17079-X (print), 3-642-17080-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
17080-5.

Capcarrere:2005:AAL

[CFB^{+}05] Mathieu S. Capcarrere, Alex A. Freitas, Peter J. Bentley, Colin G. Johnson, and Jon Timmis, editors. Advances in artificial life: 8th European conference, ECAL 2005, Canterbury, UK, September 5–9, 2005: proceedings, volume 3630 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28848-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .E87 2005. URL http:

//www.springerlink.
com/content/9783540288480;
http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3630.

Cushing:2011:SSD

[CFB11] Judith Bayard Cushing, James French, and Shawn Bowers, editors. Scientific and Statistical Database Management: 23rd International Conference, SSDBM 2011, Portland, OR, USA, July 20–22, 2011. Proceedings, volume 6809 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22350-8 (print), 3-642-22351-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22351-8.

Chouinard:1996:ITA

[CFG96] Jean-Yves Chouinard, Paul Fortier, and T. Aaron Gulliver, editors. Information theory and applications II: 4th Canadian workshop, Lac Delage, Quebec, Canada, May 28–30, 1995: Selected Papers, volume 1133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61748-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q350 .I515 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1133.htm;
 http://www.springerlink.
com/content/978-3-540-
61748-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1133.

Canadian Workshop on Information Theory.

Clough:2011:AIR

[CFG^{+}11] Paul Clough, Colum Foley, Cathal Gurrin, Gareth J. F. Jones, Wessel Kraaij, Hyowon Lee, and Vanessa Mudoch, editors. Advances in Information Retrieval: 33rd European Conference on IR Research, ECIR 2011, Dublin, Ireland, April 18–21, 2011. Proceedings, volume 6611 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20160-1 (print), 3-642-20161-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20161-5.

Catalano:2011:PKC

[CFGN11] Dario Catalano, Nelly Fazio, Rosario Gennaro, and Antonio Nicolosi, editors. Public Key Cryptography — PKC 2011: 14th International Conference on Practice and Theory in Public Key Cryptography, Taormina, Italy, March 6–9, 2011. Proceedings, volume 6571 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19378-1 (print), 3-642-19379-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19379-8.

Catarci:2012:IAE

[CFH^{+}12] Tiziana Catarci, Pamela Forner, Djoerd Hiemstra, Anselmo Peñas, and Giuseppe Santucci, editors. Information Access Evaluation. Multilinguality, Multimodality, and Visual Analytics: Third International Conference of the CLEF Initiative, CLEF 2012, Rome, Italy, September 17–20, 2012. Proceedings, volume 7488 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33246-3 (print), 3-642-33247-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33247-0.

Calamoneri:2006:ACI

[CFI06] Tiziana Calamoneri, Irene Finocchi, and Giuseppe F. Italiano, editors. Algorithms and Complexity:

6th Italian Conference, CIAC 2006, Rome, Italy, May 29–31, 2006. Proceedings, volume 3998 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34375-X (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34375-2.

Cheung:2012:HLI

[CFK^{+}12] Simon K. S. Cheung, Joseph Fong, Lam-For Kwok, Kedong Li, and Reggie Kwan, editors. Hybrid Learning: 5th International Conference, ICHL 2012, Guangzhou, China, August 13–15, 2012. Proceedings, volume 7411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32017-1 (print), 3-642-32018-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32018-7.

Croitoru:2010:CSI

[CFL10] Madalina Croitoru, Sébastien Ferré, and Dickson Lukose, editors. Conceptual Structures: From Information to Intelligence: 18th International Conference on Conceptual Structures, ICCS 2010, Kuching, Sarawak, Malaysia, July 26–30, 2010. Proceedings, volume 6208 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14196-X (print), 3-642-14197-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14197-3.

Chen:2012:FIS

[CFLR12] Li Chen, Alexander Felfernig, Jiming Liu, and Zbigniew W. Raś, editors. Foundations of Intelligent Systems: 20th International Symposium, ISMIS 2012, Macau, China, December 4–7, 2012. Proceedings, volume 7661 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34623-5 (print), 3-642-34624-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
34624-8.

Carswell:2009:WWG

[CFM09] James D. Carswell, A. Stewart Fotheringham, and Gavin McArdle, editors. Web and Wireless Geographical Information Systems: 9th International Symposium, W2GIS 2009, Maynooth, Ireland, December 7– 8, 2009. Proceedings, volume 5886 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10600-5 (print), 3-642-10601-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10601-9.

Cosnard:1994:PDC

[CFP94] M. Cosnard, Afonso Ferreira, and Joseph Peters, editors. Parallel and distributed computing: theory and practice:first Canada-France conference, Montréal, Canada, May 19–21, 1994: proceedings, volume 805 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58078-6 (New York), 3-540-58078-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P364 1994. DM58.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0805.
htm;   http://www.springerlink.
com/content/978-0-387-
58078-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=805.

Papers presented at the First Canada-France Conference on Parallel Computing.

Corne:2009:MCI

[CFP^{+}09] David Wolfe Corne, Pierluigi Frisco, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa, editors. Membrane Computing: 9th International Workshop, WMC 2008, Edinburgh, UK, July 28–31, 2008, Revised Selected and Invited Papers, volume 5391 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-95884-3 (print), 3-540-95885-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
95885-7.

Comyn:1992:LPA

[CFR92] G. Comyn, N. E. Fuchs, and M. Ratcliffe, editors. Logic programming in action: Second international Logic Programming Summer School, LPSS ’92, Zürich, Switzerland, September 7–11, 1992: proceedings, volume 636 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55930-2 (Berlin), 0-387-55930-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .L66 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0636.htm;
http://www.springerlink.
com/content/978-3-540-
55930-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=636.

Crovella:2010:NII

[CFRR10] Mark Crovella, Laura Marie Feeney, Dan Rubenstein, and S. V. Raghavan, editors. NETWORKING 2010: 9th International IFIP TC 6 Networking Conference, Chennai, India, May 11–15, 2010. Proceedings, volume 6091 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12962-5 (print), 3-642-12963-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12963-6.

Crestani:2006:SPI

[CFS06] Fabio Crestani, Paolo Ferragina, and Mark Sanderson, editors. String Processing and Information Retrieval: 13th International Conference, SPIRE 2006, Glasgow, UK, October 11– 13, 2006. Proceedings, volume 4209 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45774-7 (print), 3-540-45775-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45775-6.

Cipolla-Ficarra:2012:HCI

[CFVCF^{+}12] Francisco Cipolla-Ficarra, Kim Veltman, Huang Chih-Fang, Miguel Cipolla-Ficarra, and Andreas Kratky, editors. Human-Computer Interaction, Tourism and Cultural Heritage: Second International Workshop, HC-ITOCH 2011, Córdoba, Argentina, September 14–15, 2011, Revised Selected Papers, volume 7546 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33943-3 (print), 3-642-33944-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33944-8.

Cipolla-Ficarra:2012:ANT

[CFVV^{+}12] Francisco Cipolla-Ficarra, Kim Veltman, Domen Verber, Miguel Cipolla-Ficarra, and Florian Kammüller, editors. Advances in New Technologies, Interactive Interfaces and Communicability: Second International Conference, ADNTIIC 2011, Huerta Grande, Argentina, December 5– 7, 2011, Revised Selected Papers, volume 7547 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34009-1 (print), 3-642-34010-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34010-9.

Cao:2010:ADMb

[CFZ10] Longbing Cao, Yong Feng, and Jiang Zhong, editors. Advanced Data Mining and Applications: 6th International Conference, ADMA 2010, Chongqing, China, November 19–21, 2010, Proceedings, Part I, volume 6440 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17315-2 (print), 3-642-17316-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17316-5.

Cohen:1988:CTA

[CG88] G. Cohen and P. Godlewski, editors. Coding theory and applications: 2nd international colloquium, Cachan-Paris, France, November 24–26, 1986: proceedings, volume 311 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19368-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .C661 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0311.htm;
 http://www.springerlink.
com/content/978-0-387-
19368-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=311.

Proceedings of the colloquium “Trois journées sur le codage.”.

Crochemore:1994:CPM

[CG94] Maxime Crochemore and Dan Gusfield, editors. Combinatorial pattern matching: 5th annual symposium, CPM 94, Asilomar, CA, USA, June 5–8, 1994: proceedings, volume 807 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58094-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 1994. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0807.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        58094-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=807.

Chan:2002:ISI

[CG02] Agnes Hui Chan and Virgil D. Gligor, editors. Information security: 5th International Conference, ISC 2002, Sao Paulo, Brazil, September 30–October 2, 2002: Proceedings, volume 2433 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44270-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I89 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2433.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        44270-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2433.

Also available via the

World Wide Web.

Calzarossa:2004:PTA

[CG04] Maria Carla Calzarossa and Erol Gelenbe, editors. Performance Tools and Applications to Networked Systems: Revised Tutorial Lectures, volume 2965 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21945-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2965.htm;
 http://www.springerlink.
com/content/978-3-540-
21945-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2965;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97129.

Calder:2007:CMS

[CG07] Muffy Calder and Stephen Gilmore, editors. Computational Methods in Systems Biology: International Conference CMSB 2007, Edinburgh, Scotland, September 20– 21, 2007. Proceedings, volume 4695 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75139-4 (print), 3-540-75140-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75140-3.

Clavier:2009:CHE

[CG09] Christophe Clavier and Kris Gaj, editors. Cryptographic Hardware and Embedded Systems — CHES 2009: 11th International Workshop Lausanne, Switzerland, September 6–9, 2009 Proceedings, volume 5747 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04137-X (print), 3-642-04138-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04138-9.

Calinescu:2012:LSC

[CG12] Radu Calinescu and David Garlan, editors. Large-Scale Complex IT Systems. Development, Operation and Management: 17th Monterey Workshop 2012, Oxford, UK, March 19–21, 2012, Revised Selected Papers, volume 7539 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34058-X (print), 3-642-34059-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34059-8.

Crnkovic:2011:SAE

[CGB11] Ivica Crnković, Volker Gruhn, and Matthias Book, editors. Software Architecture: 5th European Conference, ECSA 2011, Essen, Germany, September 13–16, 2011. Proceedings, volume 6903 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23797-5 (print), 3-642-23798-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23798-0.

Cuadrado-Gallego:2008:SPP

[CGBDA08] Juan J. Cuadrado-Gallego, René Braungarten, Reiner R. Dumke, and Alain Abran, editors. Software Process and Product Measurement: International Conference, IWSM-Mensura 2007, Palma de Mallorca, Spain, November 5–8, 2007. Revised Papers, volume 4895 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85552-1 (print), 3-540-85553-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85553-8.

Chan:2008:RSC

[CGBZ08] Chien-Chung Chan, Jerzy W. Grzymala-Busse, and Wojciech P. Ziarko, editors. Rough Sets and Current Trends in Computing: 6th International Conference, RSCTC 2008 Akron, OH, USA, October 23– 25, 2008 Proceedings, volume 5306 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88423-8 (print), 3-540-88425-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88425-5.

Choren:2007:SEM

[CGG^{+}07] Ricardo Choren, Alessandro Garcia, Holger Giese, Ho fung Leung, Carlos Lucena, and Alexander Romanovsky, editors. Software Engineering for Multi-Agent Systems V: Research Issues and Practical Applications, volume 4408 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73130-X (print), 3-540-73131-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73131-3.

Conti:2003:PWC

[CGGO03] Marco Conti, Silvia Giordano, Enrico Gregori, and Stephan Olariu, editors. Personal Wireless Communications: IFIP-TC6 8th International Conference, PWC 2003, Venice, Italy, September 23–25, 2003: Proceedings, volume 2775 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20123-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.485 .I346 2003.

URL http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t2775.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20123-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2775;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12004.

Cerna:2011:STP

[ČGH^{+}11] Ivana Černá, Tibor Gyimóthy, Juraj Hromkovič, Keith Jefferey, Rastislav Králović, Marko Vukolić, and Stefan Wolf, editors. SOFSEM 2011: Theory and Practice of Computer Science: 37th Conference on Current Trends in Theory and Practice of Computer Science, Nový Smokovec, Slovakia, January 22– 28, 2011. Proceedings, volume 6543 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18380-8 (print), 3-642-18381-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18381-2.

Campos:2011:HCIa

[CGJ^{+}11a] Pedro Campos, Nicholas Graham, Joaquim Jorge, Nuno Nunes, Philippe Palanque, and Marco Winckler, editors. Human-Computer Interaction — INTERACT 2011: 13th IFIP TC 13 International Conference, Lisbon, Portugal, September 5–9, 2011, Proceedings, Part I, volume 6946 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23773-8 (print), 3-642-23774-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23774-4.

Campos:2011:HCIb

[CGJ^{+}11b] Pedro Campos, Nicholas Graham, Joaquim Jorge, Nuno Nunes, Philippe Palanque, and Marco Winckler, editors. Human-Computer Interaction — INTERACT 2011: 13th IFIP TC 13 International Conference, Lisbon, Portugal, September 5–9, 2011, Proceedings, Part II, volume 6947 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23770-3 (print), 3-642-23771-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23771-3.

Campos:2011:HCIc

[CGJ^{+}11c] Pedro Campos, Nicholas Graham, Joaquim Jorge, Nuno Nunes, Philippe Palanque, and Marco Winckler, editors. Human-Computer Interaction — INTERACT 2011: 13th IFIP TC 13 International Conference, Lisbon, Portugal, September 5–9, 2011, Proceedings, Part III, volume 6948 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23764-9 (print), 3-642-23765-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23765-2.

Campos:2011:HCId

[CGJ^{+}11d] Pedro Campos, Nicholas Graham, Joaquim Jorge, Nuno Nunes, Philippe Palanque, and Marco Winckler, editors. Human-Computer Interaction — INTERACT 2011: 13th IFIP TC 13 International Conference, Lisbon, Portugal, September 5–9, 2011, Proceedings, Part IV, volume 6949 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23767-3 (print), 3-642-23768-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23768-3.

Coello:2012:AIS

[CGK^{+}12] Carlos A. Coello Coello, Julie Greensmith, Natalio Krasnogor, Pietro Liò, Giuseppe Nicosia, and Mario Pavone, editors. Artificial Immune Systems: 11th International Conference, ICARIS 2012, Taormina, Italy, August 28–31, 2012. Proceedings, volume 7597 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33756-2 (print), 3-642-33757-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33757-4.

Chapman:2011:OPE

[CGKM11] Barbara M. Chapman, William D. Gropp, Kalyan Kumaran, and Matthias S. Müller, editors. OpenMP in the Petascale Era: 7th International Workshop on OpenMP, IWOMP 2011, Chicago, IL, USA, June 13–15, 2011. Proceedings, volume 6665 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21486-X (print), 3-642-21487-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21487-5.

Cao:2009:ADM

[CGL^{+}09] Longbing Cao, Vladimir Gorodetsky, Jiming Liu, Gerhard Weiss, and Philip S. Yu, editors. Agents and Data Mining Interaction: 4th International Workshop, ADMI 2009, Budapest, Hungary, May 10–15,2009, Revised Selected Papers, volume 5680 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03602-3 (print), 3-642-03603-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03603-3.

Chee:2011:CCT

[CGL^{+}11] Yeow Meng Chee, Zhenbo Guo, San Ling, Fengjing Shao, Yuansheng Tang, Huaxiong Wang, and Chaoping Xing, editors. Coding and Cryptology: Third International Workshop, IWCC 2011, Qingdao, China, May 30–June 3, 2011. Proceedings, volume 6639 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20900-9 (print), 3-642-20901-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20901-7.

Choren:2005:SEM

[CGLR05] Ricardo Choren, Alessandro Garcia, Carlos Lucena, and Alexander Romanovsky, editors. Software engineering for multi-agent systems III: research issues and practical applications, volume 3390 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24843-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S453. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3390;
 http://www.springerlink.
com/content/978-3-540-
24843-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3390;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106347.

Cohen:1995:AAA

[CGM95] G. Cohen, Marc Giusti, and Teo Mora, editors. Applied algebra, algebraic algorithms, and errorcorrecting codes: 11th International Symposium, AAECC-11, Paris, France, July 1995: proceedings, volume 948 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60114-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.948. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0948.htm;
http://www.springerlink.
com/content/978-3-540-
60114-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=948.

Casares-Giner:2011:NWI

[CGMP11] Vicente Casares-Giner, Pietro Manzoni, and Ana Pont, editors. NET-WORKING 2011 Workshops: International IFIP TC 6 Workshops, PE-CRN, NC-Pro, WCNS, and SUNSET 2011, Held at NETWORK-ING 2011, Valencia, Spain, May 13, 2011, Revised Selected Papers, volume 6827 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23040-7 (print), 3-642-23041-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23041-7.

Calmet:2008:MMC

[CGMQ08] Jacques Calmet, Willi Geiselmann, and Jörn Müller-Quade, editors. Mathematical Methods in Computer Science: Essays in Memory of Thomas Beth, volume 5393 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89993-6 (print), 3-540-89994-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89994-5.

Cerri:2002:ITS

[CGP02] Stefano A. Cerri, Guy Gouardères, and Fábio Paraguaçu, editors. Intelligent Tutoring Systems: 6th International Conference, ITS 2002, Biarritz, France and San Sebastian, Spain, June 2–7, 2002. Proceedings, volume 2363 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43750-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.73 .I58 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2363.htm;
 http://www.springerlink.

com/content/978-3-540-
43750-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2363.

Cimato:2003:SCN

[CGP03] Stelvio Cimato, Clemente Galdi, and Guiseppe Persiano, editors. Security in communication networks: Third International Conference, SCN 2002, Amalfi, Italy, September 11– 13, 2002: Revised Papers, volume 2576 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00420-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.59 .S385 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2576.htm;
http://www.springerlink.
com/content/978-3-540-
00420-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2576.

Also available via the World Wide Web.

Casati:2001:TSS

[CGS01] Fabio Casati, Dimitrios Georgakopoulos, and Ming-Chien Shan, editors. Technologies for E-Services: Second International Workshop, TES 2001, Rome, Italy, September 14–15, 2001: Proceedings, volume 2193 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42565-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .T43 2001; QA267.A1 L43 no.2193. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2193.htm;
 http://www.springerlink.
com/content/978-3-540-
42565-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2193.

Colosimo:2002:MDA

[CGS02] Alfredo Colosimo, Alessandro Giuliani, and Paolo Sirabella, editors. Medical Data Analysis: Third International Symposium, ISMDA 2002, Rome, Italy, October 10–11, 2002. Proceedings, volume 2526 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00044-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R853.S7 I58 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2526.htm;
http://www.springerlink.
com/content/978-3-540-
00044-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2526.

Clemm:2007:MVN

[CGS07] Alexander Clemm, Lisandro Zambenedetti Granville, and Rolf Stadler, editors. Managing Virtualization of Networks and Services: 18th IFIP/IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2007, San José, CA, USA, October 29–31, 2007. Proceedings, volume 4785 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75693-0 (print), 3-540-75694-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75694-1.

Crestani:2002:AIR

[CGvR02] Fabio Crestani, Mark Girolami, and Cornelis Joost van Rijsbergern, editors. Advances in information retrieval: 24th BCS-IRSG European Colloquium on IR Research, Glasgow, UK, March 24–27, 2002: Proceedings, volume 2291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43343-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B429 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2291.
htm;   http://www.springerlink.
com/content/978-3-540-
43343-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2291.

Colchester:1991:IPM

[CH91] A. C. F. Colchester and D. J. Hawkes, editors. Information processing in medical imaging: 12th International Conference, IPMI ’91, Wye, UK, July 7–12, 1991, proceedings, volume 511 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54246-9 (Berlin), 0-387-54246-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN RC78.7.D53 I54 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0511.htm;
 http://www.springerlink.
com/content/978-3-540-
54246-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=511.

Ciancarini:1996:CLM

[CH96] Paolo Ciancarini and Chris Hankin, editors. Coordination languages and models: First International Conference CO-ORDINATION ’96, Cesena, Italy, April 15– 17, 1996: proceedings, volume 1061 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61052-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I52 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1061.htm;
 http://www.springerlink.
com/content/978-3-540-
61052-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1061.

COORDINATION ’96.

Cluet:1998:DPL

[CH98] Sophie Cluet and Richard Hull, editors. Database programming languages: 6th International Workshop, DBPL-6, Estes Park, Colorado, USA, August 18–20, 1997: Proceedings, volume 1369 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64823-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D155925 1998.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1369.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        64823-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1369.

Christensen:1999:CVS

[CH99] H. I. (Henrik I.) Christensen and Otto Harrassowitz, editors. Computer vision systems: First International Conference, ICVS ’99, Las Palmas, Gran Canaria, Spain, January 13–15, 1999: Proceedings, volume 1542 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65459-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .I58 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1542.htm;
 http://www.springerlink.
com/content/978-3-540-
65459-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1542.

Crowcroft:2001:NGC

[CH01] Jon Crowcroft and Markus Hofmann, editors. Networked Group Communication: Third International COST264 Workshop, NGC 2001, London, UK, November 7– 9, 2001. Proceedings, volume 2233 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42824-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.887 .I58 2001; QA267.A1 L43 no.2233. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2233.htm;
 http://www.springerlink.
com/content/978-3-540-
42824-4;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2233.

Chen:2004:RTE

[CH04] Jing Chen and Seongsoo Hong, editors. Real-Time and Embedded Computing Systems and Applications: 9th International Conference, RTCSA 2003, Tainan City, Taiwan, ROC, February 18–20, 2003: Revised Papers, volume 2968 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21974-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54 .I557 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2968.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21974-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2968;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97784.

Chockler:2009:HSV

[CH09] Hana Chockler and Alan J. Hu, editors. Hardware and Software: Verification and Testing:

4th International Haifa Verification Conference, HVC 2008, Haifa, Israel, October 27–30, 2008. Proceedings, volume 5394 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01701-0 (print), 3-642-01702-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01702-5.

Chatterjee:2010:FMA

[CH10] Krishnendu Chatterjee and Thomas A. Henzinger, editors. Formal Modeling and Analysis of Timed Systems: 8th International Conference, FORMATS 2010, Klosterneuburg, Austria, September 8–10, 2010. Proceedings, volume 6246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15296-1 (print), 3-642-15297-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15297-9.

Chaghaghi:1985:TSP

[Cha85] François S. Chaghaghi. Time series package (TSPack), volume 187 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CODEN LNCSD9. ISBN 0-387-15202-4 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). 305 pp. LCCN QA280 .C39 1985. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0187.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        15202-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=187.

Chang:2002:AIS

[Cha02a] Whie Chang, editor. Advanced Internet Services and Applications: First International Workshop, AISA 2002, Seoul, Korea, August 1–2, 2002. Proceedings, volume 2402 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43968-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 A34 2002. URL http:

                                        //link.springer-ny.

com/link/service/series/
0558/tocs/t2402.htm;
http://www.springerlink.
com/content/978-3-540-
43968-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2402.

Chastek:2002:SPL

[Cha02b] Gary J. Chastek, editor. Software Product Lines: Second International Conference, SPLC 2, San Diego, CA, USA, August 19–22, 2002. Proceedings, volume 2379 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43985-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 S69 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2379.htm;
http://www.springerlink.
com/content/978-3-540-
43985-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2379.

Chaudhri:2003:WWS

[Cha03a] Akmal B. Chaudhri, editor. Web, Web-services, and database systems: NODe 2002 Web- and database-related workshops, Erfurt, Germany, October 7–10, 2002: Revised Papers, volume 2593 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00745-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .N65 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2593.htm;
 http://www.springerlink.
com/content/978-3-540-
00745-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2593.

Also available via the World Wide Web.

Chung:2003:DAA

[CHA03b] P. W. H. Chung, Chris Hinde, and Moonis Ali, editors. Developments in applied artificial intelligence: 16th International Conference on Industrial and Engineering Applications of Artificial Intelligence and Expert Systems, IEA/ AIE 2003, Loughborough, UK, June 23–26, 2003: Proceedings, volume 2718 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40455-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I556 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2718.htm;
 http://www.springerlink.
com/content/978-3-540-
40455-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2718.

Chakraborty:2005:DCI

[Cha05a] Goutam Chakraborty, editor. Distributed computing and Internet technology: second International Conference, ICD-CIT 2005, Bhubaneswar, India, December 22– 24, 2005: proceedings, volume 3816 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30999-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I25 2005.

URL http://www.springerlink.
com/content/978-3-540-
30999-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3816.

Chapman:2005:SMP

[Cha05b] Barbara M. Chapman, editor. Shared memory parallel programming with OpenMP: 5th International Workshop on OpenMP Applications and Tools, WOM-PAT 2004, Houston, TX, USA, May 17–18, 2004: Revised selected papers, volume 3349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24560-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .A1 L42 NO.3349. URL

                                        http://www.springerlink.
                                        com/content/978-3-540-
                                        24560-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3349;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105895.

Chaudron:2009:MSE

[Cha09] Michel R. V. Chaudron, editor. Models in Software Engineering: Workshops and Symposia at MODELS 2008, Toulouse, France, September 28 — October 3, 2008. Reports and Revised Selected Papers, volume 5421 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01647-2 (print), 3-642-01648-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01648-6.

Christiansen:2004:FQA

[CHAL04] Henning Christiansen, Mohand-Said Hacid, Troels Andreasen, and Henrik Legind Larsen, editors. Flexible Query Answering Systems: 6th International Conference, FQAS 2004, Lyon, France, June 24–26, 2004. Proceedings, volume 3055 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22160-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 A39 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3055.
htm;   http://www.springerlink.
com/content/978-3-540-
22160-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3055;

                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98090.

Cao:2012:NFA

[CHB^{+}12] Longbing Cao, Joshua Zhexue Huang, James Bailey, Yun Sing Koh, and Jun Luo, editors. New Frontiers in Applied Data Mining: PAKDD 2011 International Workshops, Shenzhen, China, May 24–27, 2011, Revised Selected Papers, volume 7104 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28319-5 (print), 3-642-28320-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28320-8.

Chien:2009:NGA

[CHCA09] Been-Chian Chien, Tzung-Pei Hong, Shyi-Ming Chen, and Moonis Ali, editors. Next-Generation Applied Intelligence: 22nd International Conference on Industrial, Engineering and Other Applications of Applied Intelligent Systems, IEA/AIE 2009, Tainan, Taiwan, June 24–27, 2009. Proceedings, volume 5579 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02567-6 (print), 3-642-02568-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02568-6.

Chang:2011:ETE

[CHCM11] Maiga Chang, Wu-Yuin Hwang, Ming-Puu Chen, and Wolfgang Müller, editors. Edutainment Technologies. Educational Games and Virtual Reality/Augmented Reality Applications: 6th International Conference on E-learning and Games, Edutainment 2011, Taipei, Taiwan, September 2011. Proceedings, volume 6872 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23455-0 (print), 3-642-23456-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23456-9.

Clack:1998:IFL

[CHD98] Chris Clack, Kevin Hammond, and Tony Davie, editors. Implementation of functional languages: 9th International Workshop, IFL’97, St. Andrews, Scotland, UK, September 10–12, 1997: Selected Papers, volume 1467 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64849-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1467. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1467.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64849-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1467.

Cappello:2007:RAP

[CHD07] Franck Cappello, Thomas Herault, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 14th European PVM/MPI User’s Group Meeting, Paris, France, September 30 — October 3, 2007. Proceedings, volume 4757 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75415-6 (print), 3-540-75416-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75416-9.

Cheese:1992:PEP

[Che92] A. (Andrew) Cheese.

Parallel execution of Parlog, volume 586 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55382-7 (Berlin), 0-387-55382-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 184 pp. LCCN QA76.73.P194 C44 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0586.
htm;   http://www.springerlink.
com/content/978-3-540-
55382-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=586.

Chen:1999:ACM

[Che99a] P. P. Chen, editor.

Advances in conceptual modeling: ER’99 Workshops on Evolution and Change in Data Management, Reverse Engineering in Information Systems, and the World Wide Web and Conceptual Modeling, Paris, France, November 15– 18, 1999: Proceedings, volume 1727 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66653-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 A3447 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1727.htm;
 http://www.springerlink.
com/content/978-3-540-
66653-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1727.

Chen:1999:CMC

[Che99b] Peter P. S. Chen, editor. Conceptual modeling: current issues and future directions, volume 1565 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65926-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.9.C65 C6565 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1565.htm;
 http://www.springerlink.
com/content/978-3-540-
65926-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1565.

Contains a collection of selected papers presented at the Symposium on Conceptual Modeling which was held in Los Angeles, California, on December 2, 1997.

Chen:2003:ISI

[Che03a] Hsinchun Chen, editor.

Intelligence and security informatics: first NSF/NIJ symposium, ISI 2003, Tucson, AZ, USA, June 2–3, 2003: Proceedings, volume 2665 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40189-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I86 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2665.htm;
 http://www.springerlink.
com/content/978-3-540-
40189-6;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2665.

Also available via the

World Wide Web.

Chen:2003:MDM

[Che03b] Ming-Syan Chen, editor. Mobile data management: 4th International Conference, MDM 2003, Melbourne, Australia, January 21–24, 2003: Proceedings, volume 2574 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00393-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.59 .I58 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2574.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00393-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2574.

Also available via the

World Wide Web.

Chen:2010:AAI

[Che10] Bo Chen, editor. Algorithmic Aspects in Information and Management: 6th International Conference, AAIM 2010, Weihai, China, July 19– 21, 2010. Proceedings, volume 6124 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14354-7 (print), 3-642-14355-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14355-7.

Chen:2011:CCI

[Che11] Liqun Chen, editor.

Cryptography and Coding: 13th IMA International Conference, IMACC 2011, Oxford, UK, December 12–15. Proceedings, volume 7089 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25515-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25515-1.

Chetouani:2007:ANS

[CHG^{+}07] Mohamed Chetouani, Amir Hussain, Bruno Gas, Maurice Milgram, and Jean-Luc Zarader, editors. Advances in Nonlinear Speech Processing: International Conference on Non-Linear Speech Processing, NOLISP 2007 Paris, France, May 22–25, 2007 Revised Selected Papers, volume 4885 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77346-0 (print), 3-540-77347-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77347-4.

Childs:1979:CBV

[Chi79] B. Childs, editor. Codes for boundary-value problems in ordinary differential equations: proceedings of a working conference, May 14–17, 1978, volume 76 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CODEN LNCSD9. ISBN 0-387-09554-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA372 .W76 1978. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0076.htm;
 http://www.springerlink.
com/content/978-0-387-
09554-7;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=76.

Chin:1995:IAA

[Chi95] Roland T. Chin, editor. Image analysis applications and computer graphics: Third International Computer Science Conference, ICSC ’95, Hong Kong, December 11–13, 1995: proceedings, volume 1024 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60697-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .I515 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1024.htm;
http://www.springerlink.
com/content/978-3-540-
60697-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1024.

Chittaro:2003:HCI

[Chi03] Luca Chittaro, editor.

Human-Computer Interaction with Mobile Devices and Services: 5th International Symposium, Mobile HCI 2003, Udine, Italy, September 8–11, 2003: Proceedings, volume 2795 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40821-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.59.M67 2003. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2795.
htm;   http://www.springerlink.
com/content/978-3-540-
40821-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2795;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12029.

Chin:2004:PLS

[Chi04] Wei-Ngan Chin, editor. Programming Languages and Systems: Second Asian Symposium, APLAS 2004, Taipei, Taiwan, November 4– 6, 2004, Proceedings, volume 3302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23724-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3302.htm;
 http://www.springerlink.
com/content/978-3-540-

23724-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3302;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b102225.

Correa:2006:LTI

[CHK06] José R. Correa, Alejandro Hevia, and Marcos Kiwi, editors. LATIN 2006: Theoretical Informatics: 7th Latin American Symposium, Valdivia, Chile, March 20– 24, 2006. Proceedings, volume 3887 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32755-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32755-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3887.

Cai:2010:ACI

[CHKL10] Zhihua Cai, Chengyu Hu, Zhuo Kang, and Yong Liu, editors. Advances in Computation and Intelligence: 5th International Symposium, ISICA 2010, Wuhan, China, October 22–24, 2010. Proceedings, volume 6382 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16492-7 (print), 3-642-16493-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16493-4.

Calude:2010:UCI

[CHM^{+}10] Cristian S. Calude, Masami Hagiya, Kenichi Morita, Grzegorz Rozenberg, and Jon Timmis, editors. Unconventional Computation: 9th International Conference, US 2010, Tokyo, Japan, June 21– 25, 2010. Proceedings, volume 6079 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13522-6 (print), 3-642-13523-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13523-1.

Choffrut:1988:ANL

[Cho88] Christian Choffrut, editor. Automata networks: LITP Spring School on theoretical computer science, Argeles-village, France, May 12–16, 1986: proceedings, volume 316 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19444-4 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.C45 L31 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0316.htm;
http://www.springerlink.
com/content/978-0-387-
19444-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=316.

Chong:2002:INIa

[Cho02a] Ilyoung Chong, editor.

Information networking: International Conference, ICOIN 2002, Cheju Island, Korea, January 30–February 1, 2002: revised papers, Part I. Wired Communications and Management, volume 2343 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44256-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I57145 2002 pt.1-2 (2002). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2343.
htm;   http://www.springerlink.
com/content/978-3-540-
44256-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2343.

Chong:2002:INIb

[Cho02b] Ilyoung Chong, editor. Information networking: International Conference, ICOIN 2002, Cheju Island, Korea, January 30–February 1, 2002: revised papers, Part II. Wireless Communications Technologies and Network Applications, volume 2344 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44255-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I57145 2002 pt.1-2 (2002). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2344.
htm;   http://www.springerlink.

com/content/978-3-540-
44255-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2344.

Chor:2012:RCM

[Cho12] Benny Chor, editor. Research in Computational Molecular Biology: 16th Annual International Conference, RECOMB 2012, Barcelona, Spain, April 21–24, 2012. Proceedings, volume 7262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29626-2 (print), 3-642-29627-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29627-7.

Hui:2007:TLD

[cHPkC^{+}07] Kin chuen Hui, Zhigeng Pan, Ronald Chi kit Chung, Charlie C. L. Wang, Xiaogang Jin, Stefan Göbel, and Eric C.- L. Li, editors. Technologies for E-Learning and Digital Entertainment: Second International Conference, Edutainment 2007, Hong Kong, China, June 11– 13, 2007. Proceedings, volume 4469 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73010-9 (print), 3-540-73011-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73011-8.

Castelluccia:2005:SAH

[CHPW05] Claude Castelluccia, Hannes Hartenstein, Christof Paar, and Dirk Westhoff, editors. Security in ad-hoc and sensor networks: First European workshop, ESAS 2004, Heidelberg, Germany, August 6, 2004, revised selected papers, volume 3313 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24396-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5102.85 .E97 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24396-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3313;
 http://www.springerlink.
com/openurl.asp?genre=

volume&id=doi:10.1007/
b105219.

Christodoulakis:1991:ACA

[Chr91] Dimitrios Christodoulakis, editor. Ada, the choice for ’92: Ada-Europe International Conference, Athens, Greece, May 13–17, 1991: proceedings, volume 499 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-54092-X (USA), 3-540-54092-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 A24 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0499.
htm;   http://www.springerlink.
com/content/978-0-387-
54092-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=499.

Christianson:1999:SPI

[Chr99] Bruce Christianson, editor. Security protocols: 6th International Workshop, Cambridge, UK, April 15–17, 1998: Proceedings, volume 1550 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65663-4 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 S45 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1550.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65663-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1550.

Christianson:2000:SPI

[Chr00a] Bruce Christianson, editor. Security protocols: 7th International Workshop, Cambridge, UK, April 19–21, 1999: Proceedings, volume 1796 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67381-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1796. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1796.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67381-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1796.

Christodoulakis:2000:NLP

[Chr00b] Dimitrios Christodoulakis, editor. Natural language processing—NLP 2000: Second International Conference, Patras, Greece, June 2– 4, 2000: Proceedings, volume 1835 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67605-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 I56 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1835.htm;
 http://www.springerlink.
com/content/978-3-540-
67605-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1835.

Christianson:2001:SPI

[Chr01] Bruce Christianson, editor. Security protocols: 8th International Workshop, Cambridge, UK, April 3–5, 2000: Revised Papers, volume 2133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42566-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2133. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2133.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42566-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2133.

Cunningham:2010:AMR

[CHR10] Hamish Cunningham, Allan Hanbury, and Stefan Rüger, editors. Advances in Multidisciplinary Retrieval: First Information Retrieval Facility Conference, IRFC 2010, Vienna, Austria, May 31, 2010. Proceedings, volume 6107 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13083-6 (print), 3-642-13084-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13084-7.

Chung:2003:WCT

[Chu03] Chin-Wan Chung, editor. Web and communication technologies and Internet-related social issues — HSI 2003: Second International Conference on Human.Society@Internet, Seoul, Korea, June 18– 20, 2003: proceedings, volume 2713 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40456-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 I5345 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2713.htm;
 http://www.springerlink.
com/content/978-3-540-
40456-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2713.

Chitil:2008:IAF

[CHZ08] Olaf Chitil, Zoltán Horváth, and Viktória Zsók, editors. Implementation and Application of Functional Languages: 19th International Workshop, IFL 2007, Freiburg, Germany, September 27–29, 2007. Revised Selected Papers, volume 5083 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85372-3 (print), 3-540-85373-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85373-2.

Chwa:1998:ACI

[CI98] Kyung-Yong Chwa and Oscar H. Ibarra, editors. Algorithms and Computation: 9th International Symposium, ISAAC’98, Taejon, Korea, December 14–16, 1998. Proceedings, volume 1533 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65385-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1533; QA76.9.A43 I73 1998.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1533.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65385-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1533.

Caires:2005:ALP

[CIM^{+}05] Luís Caires, Giuseppe F. Italiano, Luís Monteiro, Catuscia Palamidessi, and Moti Yung, editors. Automata, languages and programming: 32nd international colloquim, ICALP 2005, Lisbon, Portugal, July 11–15, 2005: proceedings, volume 3580 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27580-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I551 2005; QA267 .I55 2005eb; QA267 .I55 2005; QA267; Internet; QA267 .I45 2005. URL

http://www.springerlink.
com/content/978-3-540-
27580-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3580.

Cipolla:1995:AVI

[Cip95] Roberto Cipolla. Active visual inference of surface shape, volume 1016 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60642-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 193 pp. LCCN TA1634 .C57 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1016.
htm;   http://www.springerlink.
com/content/978-3-540-
60642-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1016.

Catania:2011:ADI

[CIT11] Barbara Catania, Mirjana Ivanović, and Bernhard Thalheim, editors. Advances in Databases and Information Systems: 14th East European Conference, ADBIS 2010, Novi Sad, Serbia, September 20–24, 2010. Proceedings, volume 6295 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-15575-8 (print), 3-642-15576-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15576-5.

Calmet:2006:AIS

[CIW06] Jacques Calmet, Tetsuo Ida, and Dongming Wang, editors. Artificial Intelligence and Symbolic Computation: 8th International Conference, AISC 2006 Beijing, China, September 20–22, 2006 Proceedings, volume 4120 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39728-0 (print), 3-540-39730-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
39730-4.

Choffrut:1991:S

[CJ91] Christian Choffrut and Matthias Jantzen, editors. STACS 91, volume 480 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-53709-0 (Berlin), 0-387-53709-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.480. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0480.htm;
http://www.springerlink.
com/content/978-3-540-
53709-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=480.

Comon:1995:TRF

[CJ95] Hubert Comon and Jean-Pierre Jouannaud, editors. Term rewriting: French Spring School of Theoretical Computer Science, Font Romeux [sic], France, May 17– 21, 1993: advanced course, volume 909 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59340-3 (Berlin), 0-387-59340-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .F74 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0909.htm;
 http://www.springerlink.
com/content/978-3-540-
59340-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=909.

Cho:2005:TAH

[CJ05] Kenjiro Cho and Philippe Jacquet, editors. Technologies for advanced heterogeneous networks: First Asian Internet Engineering Conference, AINTEC 2005, Bangkok, Thailand, December 13– 15, 2005: proceedings, volume 3837 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30884-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .A85 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30884-3;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3837.

Cho:2006:TAH

[CJ06] Kenjiro Cho and Philippe Jacquet, editors. Technologies for Advanced Heterogeneous Networks II: Second Asian Internet Engineering Conference, AINTEC 2006, Pathumthani, Thailand, November 28–30, 2006. Proceedings, volume 4311 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49364-6 (print), 3-540-49365-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49365-5.

Cassez:2008:FMA

[CJ08a] Franck Cassez and Claude Jard, editors. Formal Modeling and Analysis of Timed Systems: 6th International Conference, FORMATS 2008, Saint Malo, France, September 15–17, 2008. Proceedings, volume 5215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85777-X (print), 3-540-85778-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85778-5.

Chen:2008:AGM

[CJ08b] Falai Chen and Bert Jüttler, editors. Advances in Geometric Modeling and Processing: 5th International Conference, GMP 2008, Hangzhou, China, April 23–25, 2008. Proceedings, volume 4975 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79245-7 (print), 3-540-79246-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79246-8.

Calinescu:2011:FCS

[CJ11] Radu Calinescu and Ethan Jackson, editors. Foundations of Computer Software. Modeling, Development, and Verification of Adaptive Systems: 16th Monterey Workshop 2010, Redmond, WA, USA, March 31- April 2, 2010, Revised Selected Papers, volume 6662 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21291-3 (print), 3-642-21292-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21292-5.

Constable:1982:IPC

[CJE82] R. L. (Robert L.) Constable, Scott D. Johnson, and C. D. Eichenlaub. An introduction to the PL/CV2 programming logic, volume 135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CODEN LNCSD9. ISBN 0-387-11492-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 292 pp. LCCN QA76.73 .P235C75. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0135.htm;
http://www.springerlink.
com/content/978-0-387-
11492-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=135.

Chytil:1988:MFC

[CJK88] Michal Chytil, L. Janiga, and Vaclav Koubek, editors. Mathematical foundations of computer science 1988: proceedings of the 13th symposium, Carlsbad, Czechoslovakia, August 29–September 2, 1988, volume 324 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 3-540-50110-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 M3851 1988.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0324.
htm;   http://www.springerlink.
com/content/978-3-540-
50110-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=324.

Chaum:2010:TTE

[CJR^{+}10] David Chaum, Markus Jakobsson, Ronald L. Rivest, Peter Y. A. Ryan, Josh Benaloh, Miroslaw Kutylowski, and Ben Adida, editors. Towards Trustworthy Elections: New Directions in Electronic Voting, volume 6000 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12979-X (print), 3-642-12980-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12980-3.

Charikar:2007:ARC

[CJRR07] Moses Charikar, Klaus Jansen, Omer Reingold, and José D. P. Rolim, editors. Approximation, Randomization, and Combinatorial Optimization. Algorithms and Techniques: 10th International Workshop, AP-PROX 2007, and 11th International Workshop, RANDOM 2007, Princeton, NJ, USA, August 20–22, 2007. Proceedings, volume 4627 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74207-7 (print), 3-540-74208-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74208-1.

Chekuri:2005:ARC

[CJRT05] Chandra Chekuri, Klaus Jansen, José D. P. Rolim, and Luca Trevisan, editors. Approximation, randomization, and combinatorial optimization: algorithms and techniques: 8th International Workshop on Approximation Algorithms for Combinatorial Optimization Problems, APPROX 2005, and 9th International Workshop on Randomization and Computation, RANDOM 2005, Berkeley, CA, USA, August 22–24, 2005: proceedings, volume 3624 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28239-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I643 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3624;
 http://www.springerlink.
com/content/978-3-540-
28239-6.

Cremers:1982:TCS

[CK82] A. B. Cremers and H. P. Kriegel, editors. Theoretical computer science: 6th GI-conference, Dortmund, January 5– 7, 1983, volume 145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11973-6 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .T473 1982. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0145.htm;
 http://www.springerlink.
com/content/978-0-387-
11973-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=145.

Papers presented at the 6th Biannual GI Symposium on Theoretical Computer Science, held in Dortmund, Germany, Jan. 5–7, 1983, and sponsored by the Special Interest Group for Automata and Formal Language Theory of the Gesellschaft für Informatik.

Chytil:1984:MFC

[CK84] Michal Chytil and Vaclav Koubek, editors. Mathematical foundations of computer science 1984: proceedings, 11th Symposium, Praha, Czechoslovakia, September 3–7, 1984, volume 176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-13372-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 M38 1984. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0176.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        13372-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=176.

Consists of contributions presented at the 11th Symposium on Mathematical Foundations of Computer Science held at Prague, Czechoslovakia, Sept. 3–7, 1984.

Clarke:1991:CAV

[CK91] E. M. Clarke and R. P. Kurshan. Computeraided verification: proceedings, volume 531 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54477-1 (Berlin), 0-387-54477-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 372 pp. LCCN TK7874 .D553 1990a.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0531.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        54477-7;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=531.

Csetverikov:1993:CAI

[CK93] Dmitrij Csetverikov and W. Kropatsch, editors. Computer analysis of images and patterns: 5th International Conference, CAIP ’93, Budapest, Hungary, September 13–15, 1993: proceedings, volume 719 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57233-3 (Berlin), 0-387-57233-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .C656 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0719.htm;
http://www.springerlink.
com/content/978-3-540-
57233-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=719.

Colom:2001:ATP

[CK01] Jose-Manuel Colom and Maciej Koutny, editors. Applications and theory of Petri nets 2001: 22nd International Conference, ICATPN 2001, Newcastle Upon Tyne, UK, June, 25–29, 2001: Proceedings, volume 2075 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42252-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I56 2001; QA267 .A66 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2075.htm;
 http://www.springerlink.
com/content/978-3-540-
42252-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2075.

Campilho:2004:IARa

[CK04a] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: International Conference, ICIAR 2004, Porto, Portugal, September 29– October 1, 2004, Proceedings, Part I, volume 3211 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23223-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3211.htm;
 http://www.springerlink.

com/content/978-3-540-
23223-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3211;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100437.

Campilho:2004:IARb

[CK04b] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: International Conference ICIAR 2004, Porto, Portugal, September 29–October 1, 2004, Proceedings, Part II, volume 3212 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23240-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3212.htm;
http://www.springerlink.
com/content/978-3-540-
23240-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3212;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100438.

Campilho:2006:IARb

[CK06a] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: Third International Conference, ICIAR 2006, Póvoa de Varzim, Portugal, September 18–20, 2006, Proceedings, Part II, volume 4142 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44894-2 (print), 3-540-44896-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
44896-9.

Campilho:2006:IARa

[CK06b] Aurélio Campilho and Mohamed S. Kamel, editors. Image Analysis and Recognition: Third International Conference, ICIAR 2006, Póvoa de Varzim, Portugal, September 18–20, 2006, Proceedings, Part I, volume 4141 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44891-8 (print), 3-540-44893-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
44893-8.

Chong:2006:INA

[CK06c] Ilyoung Chong and Kenji Kawahara, editors. Information Networking. Advances in Data Communications and Wireless Networks: International Conference, ICOIN 2006, Sendai, Japan, January 16–19, 2006, Revised Selected Papers, volume 3961 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48563-5 (print), 3-540-48564-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48564-3.

Cooper:2007:DMD

[CK07] Richard Cooper and Jessie Kennedy, editors. Data Management. Data, Data Everywhere: 24th British National Conference on Databases, BN-COD 24, Glasgow, UK, July 3–5, 2007. Proceedings, volume 4587 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73389-2 (print), 3-540-73390-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73390-4.

Campilho:2008:IAR

[CK08] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: 5th International Conference, ICIAR 2008, Póvoa de Varzim, Portugal, June 25–27, 2008. Proceedings, volume 5112 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69811-6 (print), 3-540-69812-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69812-8.

Campilho:2010:IARa

[CK10a] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: 7th International Conference, ICIAR 2010, Póvoa de Varzim, Portugal, June 21–23, 2010. Proceedings, Part I, volume 6111 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13771-7 (print), 3-642-13772-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13772-3.

Campilho:2010:IARb

[CK10b] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: 7th International Conference, ICIAR 2010, Póvoa de Varzin, Portugal, June 21–23, 2010, Proceedings, Part II, volume 6112 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13774-1 (print), 3-642-13775-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13775-4.

Camenisch:2012:OPN

[CK12a] Jan Camenisch and Dogan Kesdogan, editors. Open Problems in Network Security: IFIP WG 11.4 International Workshop, iNetSec 2011, Lucerne, Switzerland, June 9, 2011, Revised Selected Papers, volume 7039 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27584-2 (print), 3-642-27585-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27585-2.

Campilho:2012:IARa

[CK12b] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: 9th International Conference, ICIAR 2012, Aveiro, Portugal, June 25–27, 2012. Proceedings, Part I, volume 7324 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31294-2 (print), 3-642-31295-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
31295-3.

Campilho:2012:IARb

[CK12c] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: 9th International Conference, ICIAR 2012, Aveiro, Portugal, June 25–27, 2012. Proceedings, Part II, volume 7325 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31297-7 (print), 3-642-31298-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31298-4.

Corradini:2011:ACC

[CKC11] Andrea Corradini, Bartek Klin, and Corina Cîrstea, editors. Algebra and Coalgebra in Computer Science: 4th International Conference, CALCO 2011, Winchester, UK, August 30 — September 2, 2011. Proceedings, volume 6859 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22943-3 (print), 3-642-22944-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22944-2.

Camenisch:2011:ORP

[CKD11] Jan Camenisch, Valentin Kisimov, and Maria Dubovitskaya, editors. Open Research Problems in Network Security: IFIP WG 11.4 International Workshop, iNetSec 2010, Sofia, Bulgaria, March 5–6, 2010, Revised Selected Papers, volume 6555 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19227-0 (print), 3-642-19228-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19228-9.

Chowdhury:2010:RDL

[CKH10] Gobinda Chowdhury, Chris Koo, and Jane Hunter, editors. The Role of Digital Libraries in a Time of Global Change: 12th International Conference on Asia-Pacific Digital Libraries, ICADL 2010, Gold Coast, Australia, June 21–25, 2010. Proceedings, volume 6102 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13653-2 (print), 3-642-13654-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13654-2.

Chang:2009:LPG

[CKK^{+}09] Maiga Chang, Rita Kuo, Kinshuk, Gwo-Dong Chen, and Michitaka Hirose, editors. Learning by Playing. Game-based Education System Design and Development: 4th International Conference on E-Learning and Games, Edutainment 2009, Banff, Canada, August 9–11, 2009. Proceedings, volume 5670 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03363-6 (print), 3-642-03364-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03364-3.

Cox:2005:DWT

[CKL05] I. J. (Ingemar J.) Cox, Ton Kalker, and Heung-Kyu Lee, editors. Digital watermarking: third international workshop, IWDW 2004, Seoul, Korea, October 30– November 1, 2004: revised selected papers, volume 3304 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24839-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I89 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24839-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3304.

Cuyt:2009:NVC

[CKLM09] Annie Cuyt, Walter Krämer, Wolfram Luther, and Peter Markstein, editors. Numerical Validation in Current Hardware Architectures: International Dagstuhl Seminar, Dagstuhl Castle, Germany, January 6– 11, 2008. Revised Papers, volume 5492 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-642-01591-3, 3-642-01590-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C62 N86 2009. URL http:

//springerlink.com/
openurl.asp?genre=book%
26isbn=978-3-642-01590-
8;   http://www.springerlink.
com/content/978-3-642-
01591-5;   http://www.
springerlink.com/openurl.
asp?genre=book%26isbn=
978-3-642-01590-8.

Cheng:2010:IRT

[CKLN10] Pu-Jen Cheng, Min-Yen Kan, Wai Lam, and Preslav Nakov, editors. Information Retrieval Technology: 6th Asia Information Retrieval Societies Conference, AIRS 2010, Taipei, Taiwan, December 1–3, 2010. Proceedings, volume 6458 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17186-9 (print), 3-642-17187-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17187-1.

Clarke:1993:SQA

[CKM93] Michael Clarke, Rudolf Kruse, and Serafin Moral, editors. Symbolic and quantitative approaches to reasoning and uncertainty: European Conference ECSQARU ’93, Granada, Spain, November 8–10, 1993: proceedings, volume 747 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57395-X (USA), 3-540-57395-X (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q375 .E88 1993. DM72.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0747.htm;
 http://www.springerlink.
com/content/978-0-387-
57395-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=747.

Chua:2009:SMI

[CKM^{+}09] Tat-Seng Chua, Yiannis Kompatsiaris, Bernard Mérialdo, Werner Haas, Georg Thallinger, and Werner Bailer, editors. Semantic Multimedia: 4th International Conference on Semantic and Digital Media Technologies, SAMT 2009 Graz, Austria, December 2– 4, 2009 Proceedings, volume 5887 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10542-4 (print), 3-642-10543-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10543-2.

Comaniciu:2005:SMV

[CKMS05] Dorin Comaniciu, Kenichi Kanatani, Rudolf Mester, and David Suter, editors. Statistical methods in video processing: ECCV 2004 workshop SMVP 2004, Prague, Czech Republic, May 16, 2004, revised selected papers, volume 3247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-23989-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
23989-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3247;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104157.

Chong:2001:IMS

[CKO01] Frederic T. Chong, Christoforos Kozyrakis, and Mark Oskin, editors. Intelligent memory systems: Second International Workshop, IMS 2000, Cambridge, MA, USA, November 12, 2000: Revised Papers, volume 2107 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42328-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M45 I94 2000; QA267.A1 L43 no.2107. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2107.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42328-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2107.

Cooper:2011:HAI

[CKOB11] Eric W. Cooper, Victor V. Kryssanov, Hitoshi Ogawa, and Stephen Brewster, editors. Haptic and Audio Interaction Design: 6th International Workshop, HAID 2011, Kusatsu, Japan, August 25–26, 2011. Proceedings, volume 6851 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22949-2 (print), 3-642-22950-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22950-3.

Cortier:2009:FPS

[CKOS09] Véronique Cortier, Claude Kirchner, Mitsuhiro Okada, and Hideki Sakurada, editors. Formal to Practical Security: Papers Issued from the 2005-2008 French-Japanese Collaboration, volume 5458 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02001-1 (print), 3-642-02002-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02002-5.

Camara:2012:SAC

[CKP^{+}12] Oscar Camara, Ender Konukoglu, Mihaela Pop, Kawal Rhode, Maxime Sermesant, and Alistair Young, editors. Statistical Atlases and Computational Models of the Heart. Imaging and Modelling Challenges: Second International Workshop, STACOM 2011, Held in Conjunction with MICCAI 2011, Toronto, ON, Canada, September 22, 2011, Revised Selected Papers, volume 7085 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28325-X (print), 3-642-28326-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28326-0.

Ceberio:2007:ICC

[CKPB07] M. Ceberio, V. Kreinovich, A. Pownuk, and B. Bede. From interval computations to constraintrelated set computations: Towards faster estimation of statistics and ODEs under interval, p-box, and fuzzy uncertainty. In Patricia Melin et al., editors, Foundations of Fuzzy Logic and Soft Computing, 12th International Fuzzy Systems Association World Congress, IFSA 2007, Cancun, Mexico, June 18–21, 2007. Proceedings, volume 4549 of Lecture Notes in Computer Science, page ?? Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-72917-8. LCCN

???? URL http://www.
cs.utep.edu/vladik/
2006/tr06-50.pdf;
http://www.springerlink.
com/content/978-3-540-
72917-4.

Calude:2011:UCI

[CKPR11] Cristian S. Calude, Jarkko Kari, Ion Petre, and Grzegorz Rozenberg, editors. Unconventional Computation: 10th International Conference, UC 2011, Turku, Finland, June 6–10, 2011. Proceedings, volume 6714 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21340-5 (print), 3-642-21341-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21341-0.

Camacho:2004:ILP

[CKS04] Rui Camacho, Ross King, and Ashwin Srinivasan, editors. Inductive Logic Programming: 14th International Conference, Ilp 2004, Porto, Portugal, September 6– 8, 2004: Proceedings, volume 3194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22941-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5102.9 .I32 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3194.htm;
 http://www.springerlink.
com/content/978-3-540-
22941-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3194;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b10011.

Cossentino:2012:MAS

[CKTW12]  Massimo  Cossentino,

Michael Kaisers, Karl Tuyls, and Gerhard Weiss, editors. Multi-Agent Systems: 9th European Workshop, EU-MAS 2011, Maastricht, The Netherlands, November 14–15, 2011. Revised Selected Papers, volume 7541 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34798-3 (print), 3-642-34799-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34799-3.

Creignou:2008:CCO

[CKV08] Nadia Creignou, Phokion G.

Kolaitis, and Heribert Vollmer, editors. Complexity of Constraints: An Overview of Current Research Themes, volume 5250 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92799-9 (print), 3-540-92800-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92800-3.

Corchado:2011:HAIa

[CKW11a] Emilio Corchado, Marek Kurzyński, and Michał Woźniak, editors. Hybrid Artificial Intelligent Systems: 6th International Conference, HAIS 2011, Wroclaw, Poland, May 23–25, 2011, Proceedings, Part I, volume 6678 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21218-2 (print), 3-642-21219-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21219-2.

Corchado:2011:HAIb

[CKW11b] Emilio Corchado, Marek Kurzyński, and Michał Woźniak, editors. Hybrid Artificial Intelligent Systems: 6th International Conference, HAIS 2011, Wroclaw, Poland, May 23–25, 2011, Proceedings, Part II, volume 6679 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21221-2 (print), 3-642-21222-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21222-2.

Choffrut:1990:SAS

[CL90] Christian Choffrut and T. Lengauer, editors. STACS 90: 7th Annual Symposium on Theoretical Aspects of Computer Science, Rouen, France, February 22–24, 1990: proceedings, volume 415 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52282-4 (Berlin), 0-387-52282-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.415. DM47.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0415.
htm;   http://www.springerlink.
com/content/978-3-540-
52282-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=415.

“Organized jointly by the special interest group for Applied Mathematics of AFCET (Association française de cybernetique économique et technique), and the Special Interest Group for Theoretical Computer Sciences of GI (Gesellschaft für Informatik)”–Foreword.

Calmet:1996:DIS

[CL96] Jacques Calmet and Carla Limongelli, editors. Design and implementation of symbolic computation systems: International Symposium, DISCO ’96, Karlsruhe, Germany, September 18–20, 1996: proceedings, volume 1128 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61697-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88I576 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1128.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61697-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1128.

Castelfranchi:2001:IAV

[CL01] Cristiano Castelfranchi and Yves Lespérance, editors. Intelligent agents VII: agent theories architectures and languages: 7th International Workshop, ATAL 2000, Boston, MA, USA, July 7–9, 2000: Proceedings, volume 1986 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42422-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA267.A1 L43 no.1986; Q336 .E23 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1986.htm;
 http://www.springerlink.
com/content/978-3-540-
42422-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1986.

Chi:2004:CCA

[CL04] Chi-Hung Chi and Kwok Yan Lam, editors. Content computing: Advanced Workshop on Content Computing, AWCC 2004, ZhenJiang, JiangSu, China, November 15–17, 2004: proceedings, volume 3309 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23898-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 A333 2004. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3309;   http:
//www.springerlink.
com/content/978-3-540-
23898-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3309;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b103383.

Chang:2006:AIV

[CL06a] Long-Wen Chang and Wen-Nung Lie, editors. Advances in Image and Video Technology: First Pacific Rim Symposium, PSIVT 2006, Hsinchu, Taiwan, December 10– 13, 2006. Proceedings, volume 4319 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68297-X (print), 3-540-68298-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68298-1.

Chen:2006:CCA

[CL06b] Danny Z. Chen and D. T. Lee, editors. Computing and Combinatorics: 12th Annual International Conference, COCOON 2006, Taipei, Taiwan, August 15–18, 2006. Proceedings, volume 4112 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36925-2 (print), 3-540-36926-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36926-4.

Cerin:2007:AGP

[CL07] Christophe Cérin and Kuan-Ching Li, editors. Advances in Grid and Pervasive Computing: Second International Conference, GPC 2007, Paris, France, May 2–4, 2007. Proceedings, volume 4459 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72359-5 (print), 3-540-72360-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72360-8.

Calvanese:2008:WRR

[CL08a] Diego Calvanese and Georg Lausen, editors. Web Reasoning and Rule Systems: Second International Conference, RR 2008, Karlsruhe, Germany, October 31– November 1, 2008. Proceedings, volume 5341 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88736-9 (print), 3-540-88737-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88737-9.

Crandall:2008:ABI

[CL08b] Keith A. Crandall and Jens Lagergren, editors. Algorithms in Bioinformatics: 8th International Workshop, WABI 2008, Karlsruhe, Germany, September 15– 19, 2008. Proceedings, volume 5251 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87360-0 (print), 3-540-87361-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87361-7.

Chavez:2010:SPI

[CL10] Edgar Chavez and Stefano Lonardi, editors. String Processing and Information Retrieval: 17th International Symposium, SPIRE 2010, Los Cabos, Mexico, October 11–13, 2010. Proceedings, volume 6393 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16320-3 (print), 3-642-16321-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16321-0.

Camenisch:2011:PKI

[CL11] Jan Camenisch and Costas Lambrinoudakis, editors. Public Key Infrastructures, Services and Applications: 7th European Workshop, EuroPKI 2010, Athens, Greece, September 23– 24, 2010. Revised Selected Papers, volume 6711 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22632-9 (print), 3-642-22633-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22633-5.

Claramunt:2011:GSI

[CLB11] Christophe Claramunt, Sergei Levashkin, and Michela Bertolotto, editors. GeoSpatial Semantics: 4th International Conference, GeoS 2011, Brest, France, May 12–13, 2011. Proceedings, volume 6631 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20629-8 (print), 3-642-20630-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20630-6.

Cuellar:2011:STM

[CLBP11] Jorge Cuellar, Javier Lopez, Gilles Barthe, and Alexander Pretschner, editors. Security and Trust Management: 6th International Workshop, STM 2010, Athens, Greece, September 23– 24, 2010, Revised Selected Papers, volume 6710 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22443-1 (print), 3-642-22444-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22444-7.

Cleaveland:1992:CTI

[Cle92] W. R. Cleaveland, editor.

CONCUR ’92: Third International Conference on Concurrency Theory, Stony Brook, NY, USA, August 24–27, 1992: proceedings, volume 630 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55822-5 (Berlin), 0-387-55822-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0630.htm;
http://www.springerlink.
com/content/978-3-540-
55822-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=630.

Cleaveland:1999:TAC

[Cle99] W. R. (Walter Rance)

Cleaveland, editor. Tools and algorithms for the construction and analysis of systems: 5th International Conference, TACAS’99, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS’99, Amsterdam, The Netherlands, March 22–28, 1999: Proceedings, volume 1579 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65703-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T33 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1579.htm;
 http://www.springerlink.
com/content/978-3-540-
65703-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1579.

Comon-Lundh:2007:RCP

[CLKK07] Hubert Comon-Lundh, Claude Kirchner, and Hélène Kirchner, editors. Rewriting, Computation and Proof: Essays Dedicated to Jean-Pierre Jouannaud on the Occasion of His 60th Birthday, volume 4600 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73146-6 (print), 3-540-73147-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73147-4.

Cai:2009:ACI

[CLKL09] Zhihua Cai, Zhenhua Li, Zhuo Kang, and Yong Liu, editors. Advances in Computation and Intelligence: 4th International Symposium, ISICA 2009 Huangshi, China, October 23–25, 2009 Proceedings, volume 5821 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04842-0 (print), 3-642-04843-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04843-2.

Chee:2009:CCS

[CLL^{+}09] Yeow Meng Chee, Chao Li, San Ling, Huaxiong Wang, and Chaoping Xing, editors. Coding and Cryptology: Second International Workshop, IWCC 2009, Zhangjiajie, China, June 1–5, 2009. Proceedings, volume 5557 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01813-0 (print), 3-642-01877-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01877-0.

Chen:2009:DSA

[CLLD09] Lei Chen, Chengfei Liu, Qing Liu, and Ke Deng, editors. Database Systems for Advanced Applications: DASFAA 2009 International Workshops: BenchmarX, MCIS, WDPP, PPDA, MBC, PhD, Brisbane, Australia, April 20 — 23, 2009, volume 5667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04204-X (print), 3-642-04205-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04205-8.

Calvanese:2003:DTI

[CLM03] Diego Calvanese, Maurizio Lenzerini, and Rajeev Motwani, editors. Database theory, ICDT 2003: 9th International Conference, Siena, Italy, January 8–10, 2003: Proceedings, volume 2572 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00323-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I557 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2572.htm;
 http://www.springerlink.
com/content/978-3-540-
00323-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2572.

Also available via the World Wide Web.

Callaghan:2002:TPP

[CLMP02] Paul Callaghan, Zhaohui Luo, James McKinna, and Robert Pollack, editors. Types for proofs and programs: International Workshop, TYPES 2000, Durham, UK, December 8–12, 2000: Selected Papers, volume 2277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43287-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T96 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2277.
htm;   http://www.springerlink.
com/content/978-3-540-
43287-6;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2277.

Coppo:2005:TCS

[CLP05] Mario Coppo, Elena (Elena) Lodi, and G. Michele Pinna, editors. Theoretical computer science: 9th Italian conference, ICTCS 2005, Siena, Italy, October 12–14, 2005: proceedings, volume 3701 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29106-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I28 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29106-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3701.

Coninx:2007:TMD

[CLS07a] Karin Coninx, Kris Luyten, and Kevin A. Schneider, editors. Task Models and Diagrams for Users Interface Design: 5th International Workshop, TAMODIA 2006, Hasselt, Belgium, October 23–24, 2006. Revised Papers, volume 4385 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70815-4 (print), 3-540-70816-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70816-2.

Cooper:2007:CLR

[CLS07b] S. Barry Cooper, Benedikt Löwe, and Andrea Sorbi, editors. Computation and Logic in the Real World: Third Conference on Computability in Europe, CiE 2007, Siena, Italy, June 18–23, 2007. Proceedings, volume 4497 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73000-1 (print), 3-540-73001-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73001-9.

Camenisch:2011:DPP

[CLS11] Jan Camenisch, Ronald Leenes, and Dieter Sommer, editors. Digital Privacy: PRIME — Privacy and Identity Management for Europe, volume 6545 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19049-9 (print), 3-642-19050-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19050-6.

Cooper:2005:NCP

[CLT05] S. Barry Cooper, Benedikt Löwe, and Leen Torenvliet, editors. New computational paradigms: first conference on computability in Europe, CIE 2005, Amsterdam, The Netherlands, June 8– 12, 2005, proceedings, volume 3526 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26179-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.59 .C67 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26179-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3526;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136981.

Cao:2008:NPC

[CLWC08] Jian Cao, Minglu Li, Min-You Wu, and Jinjun Chen, editors. Network and Parallel Computing: IFIP International Conference, NPC 2008, Shanghai, China, October 18–20, 2008. Proceedings, volume 5245 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88139-5 (print), 3-540-88140-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88140-7.

Chen:2002:AKD

[CLY02] Ming-Syan Chen, Bing Liu, and Philip S. Yu, editors. Advances in Knowledge Discovery and Data Mining: 6th Pacific-Asia Conference, PAKDD 2002, Taipei, Taiwan, May 6–8, 2002. Proceedings, volume 2336 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43704-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 P17 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2336.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43704-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2336.

Cheng:2006:WAS

[CLZ06] Xiuzhen Cheng, Wei Li, and Taieb Znati, editors. Wireless Algorithms, Systems, and Applications: First International Conference, WASA 2006, Xi’an, China, August 15–17, 2006. Proceedings, volume 4138 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37189-3 (print), 3-540-37190-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37190-8.

Chen:2009:AWN

[CLZ^{+}09] Lei Chen, Chengfei Liu, Xiao Zhang, Shan Wang, Darijus Strasunskas, Stein L. Tomassen, Jinghai Rao, Wen-Syan Li, K. Selçuk Candan, Dickson K. W. Chiu, Yi Zhuang, Clarence A. Ellis, and Kwang-Hoon Kim, editors. Advances in Web and Network Technologies, and Information Management: APWeb/WAIM 2009 International Workshops: WCMT 2009, RTBI 2009, DBIR-ENQOIR 2009, PAIS 2009, Suzhou, China, April 2–4, 2009, Revised Selected Papers, volume 5731 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03995-2 (print), 3-642-03996-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03996-6.

Castelfranchi:1995:RCE

[CM95] Cristiano Castelfranchi and Jean-Pierre Muller, editors. From reaction to cognition: 5th European Workshop on Modelling Autonomous Agents in a Multi-Agent World, MAAMAW ’93, Neuchatel, Switzerland, August 1993: Selected Papers, volume 957 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60155-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .E99 1993.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0957.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60155-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=957.

Connor:2000:RIS

[CM00] Richard Connor and Alberto O. Mendelzon, editors. Research Issues in Structured and Semistructured Database Programming: 7th International Workshop on Database Programming Languages, DBPL’99, Kinloch Rannoch, UK, September 1–3, 1999. Revised Papers, volume 1949 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41481-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I585 1999. URL http:

                                        //link.springer-ny.

com/link/service/series/
0558/tocs/t1949.htm;
http://www.springerlink.
com/content/978-3-540-
41481-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1949.

Champarnaud:2003:IAA

[CM03a] Jean-Marc Champarnaud and Denis Maurel, editors. Implementation and application of automata: 7th International Conference, CIAA 2002, Tours, France, July 3–5, 2002: Revised Papers, volume 2608 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40391-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I5787 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2608.htm;
http://www.springerlink.
com/content/978-3-540-
40391-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2608.

Chico:2003:ICS

[CM03b] Jorge Juan Chico and Enrico Macii, editors. Integrated Circuit and System Design: Power and timing modeling, optimization and simulation: 13th International Workshop, PATMOS 2002, Torino, Italy, September 10–12, 2003: Proceedings, volume 2799 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20074-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.75. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2799.htm;
 http://www.springerlink.
com/content/978-3-540-
20074-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2799;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12033.

Chwa:2004:CCA

[CM04a] Kyung-Yong Chwa and J. Ian Munro, editors. Computing and Combinatorics: 10th annual International Conference, COCOON 2004, Jeju Island, Korea, August 17–20, 2004: Proceedings, volume 3106 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22856-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I5443 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3106.htm;
 http://www.springerlink.
com/content/978-3-540-
22856-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3106;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99243.

Cotin:2004:MSI

[CM04b] Stéphane Cotin and Dimitris N. Metaxas, editors. Medical Simulation: International Symposium, ISMS 2004, Cambridge, MA, USA, June 17–18, 2004: Proceedings, volume 3078 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22186-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R859.7.C65 M44 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3078.
htm;   http://www.springerlink.
com/content/978-3-540-
22186-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3078;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98155.

Cohn:2005:SIT

[CM05a] Anthony G. Cohn and David M. Mark, editors. Spatial information theory: international conference, COSIT 2005, Elliottville, NY, USA, September 14–18, 2005: proceedings, volume 3693 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28964-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.212 .C68 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28964-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3693.

Cunha:2005:EPP

[CM05b] José C. Cunha and Pedro D. Medeiros, editors. EURO-PAR 2005 parallel processing: 11th international EURO-PAR conference, Lisbon, Portugal, August 30– September 2, 2005, proceedings, volume 3648 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28700-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28700-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3648.

Chung:2006:AGP

[CM06] Yeh-Ching Chung and José E. Moreira, editors. Advances in Grid and Pervasive Computing: First International Conference, GPC 2006, Taichung, Taiwan, May 3–5, 2006: Proceedings, volume 3947 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33809-8 (paperback). LCCN QA76.9.C58 G63 2006. URL http:

//www.springerlink.
com/content/978-3-540-
33809-3.

Childs:2009:TQC

[CM09a] Andrew Childs and Michele Mosca, editors. Theory of Quantum Computation, Communication, and Cryptography: 4th Workshop,TQC 2009, Waterloo, Canada, May 11–13, 2009, Revised Selected Papers, volume 5906 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10697-8 (print), 3-642-10698-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10698-9.

Ciccarelli:2009:CGI

[CM09b] Francesca D. Ciccarelli and István Miklós, editors. Comparative Genomics: International Workshop, RECOMB-CG 2009, Budapest, Hungary, September 27– 29, 2009. Proceedings, volume 5817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04743-2 (print), 3-642-04744-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04744-2.

Corradini:2009:RTA

[CM09c] Andrea Corradini and Ugo Montanari, editors. Recent Trends in Algebraic Development Techniques: 19th International Workshop, WADT 2008, Pisa, Italy, June 13–16, 2008, Revised Selected Papers, volume 5486 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03428-4 (print), 3-642-03429-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03429-9.

Cowling:2010:ECC

[CM10] Peter Cowling and Peter Merz, editors. Evolutionary Computation in Combinatorial Optimization: 10th European Conference, Evo-COP 2010, Istanbul, Turkey, April 7–9, 2010. Proceedings, volume 6022 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12138-1 (print), 3-642-12139-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12139-5.

Cousot:2011:SAI

[CM11] Radhia Cousot and Matthieu Martel, editors. Static Analysis: 17th International Symposium, SAS 2010, Perpignan, France, September 14–16, 2010. Proceedings, volume 6337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-15768-8 (print), 3-642-15769-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15769-1.

Chakraborty:2012:ATV

[CM12] Supratik Chakraborty and Madhavan Mukund, editors. Automated technology for verification and analysis: 10th International Symposium, ATVA 2012, Thiruvananthapuram, India, October 3–6, 2012. Proceedings, volume 7561 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33385-0, 3-642-33386-9 (e-book). LCCN QA76.9.A96 A88 2012.

                                        URL http://www.springerlink.

com/content/978-3-642-
33386-6.

Cova:2008:GIS

[CMB^{+}08] Thomas J. Cova, Harvey J. Miller, Kate Beard, Andrew U. Frank, and Michael F. Goodchild, editors. Geographic Information Science: 5th International Conference, GIScience 2008, Park City, UT, USA, September 23–26, 2008. Proceedings, volume 5266 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87472-0 (print), 3-540-87473-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87473-7.

Cooper:2011:LCP

[CMCS11] Keith Cooper, John Mellor-Crummey, and Vivek Sarkar, editors. Languages and Compilers for Parallel Computing: 23rd International Workshop, LCPC 2010, Houston, TX, USA, October 7–9, 2010. Revised Selected Papers, volume 6548 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19594-6 (print), 3-642-19595-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19595-2.

Collet:2010:AEI

[CML^{+}10] Pierre Collet, Nicolas Monmarché, Pierrick Legrand, Marc Schoenauer, and Evelyne Lutton, editors. Artifical Evolution: 9th International Conference, Evolution Artificielle, EA, 2009, Strasbourg, France, October 26–28, 2009. Revised Selected Papers, volume 5975 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14155-2 (print), 3-642-14156-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14156-0.

Cohen:1993:AAA

[CMM93] G. Cohen, Teo Mora, and Oscar Moreno, editors. Applied algebra, algebraic algorithms, and errorcorrecting codes: 10th International Symposium, AAECC-10, San Juan de Puerto Rico, Puerto Rico, May 10–14, 1993: proceedings, volume 673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56686-4 (Berlin), 0-387-56686-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 1993. DM72.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0673.htm;
 http://www.springerlink.
com/content/978-3-540-
56686-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=673.

Crespo:2001:MDA

[CMM01] Jose Crespo, Victor Maojo, and Fernando Martin, editors. Medical data analysis: Second International Symposium, ISMDA 2001, Madrid, Spain, October 8–9, 2001: Proceedings, volume 2199 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42734-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R853.S7 M43 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2199.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42734-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2199.

Papers from the 2nd International Symposium on Medical Data Analysis.

Chaudhury:2009:PRM

[CMM^{+}09a] Santanu Chaudhury, Sushmita Mitra, C. A. Murthy, P. S. Sastry, and Sankar K. Pal, editors. Pattern Recognition and Machine Intelligence: Third International Conference, PReMI 2009 New Delhi, India, December 16– 20, 2009 Proceedings, volume 5909 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-11163-7 (print), 3-642-11164-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11164-8.

Chen:2009:TCS

[CMM09b] Liqun Chen, Chris J.

Mitchell, and Andrew Martin, editors. Trusted Computing: Second International Conference, Trust 2009, Oxford, UK, April 6–8, 2009, Proceedings., volume 5471 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00586-1 (print), 3-642-00587-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00587-9.

Christianson:2011:SPXa

[CMMR11] Bruce Christianson, James A.

Malcolm, Vashek Matyas, and Michael Roe, editors. Security Protocols XVI: 16th International Workshop, Cambridge, UK, April 16–18, 2008. Revised Selected Papers, volume 6615 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22136-X (print), 3-642-22137-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22137-8.

Chapman:2012:OHW

[CMMR12] Barbara M. Chapman, Federico Massaioli, Matthias S. Müller, and Marco Rorro, editors. OpenMP in a Heterogeneous World: 8th International Workshop on OpenMP, IWOMP 2012, Rome, Italy, June 11–13, 2012. Proceedings, volume 7312 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30960-7 (print), 3-642-30961-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30961-8.

Cremers:2011:VPC

[CMOZM11] Daniel Cremers, Marcus Magnor, Martin R. Oswald, and Lihi Zelnik-Manor, editors. Video Processing and Computational Video: International Seminar, Dagstuhl Castle, Germany, October 10–15, 2010. Revised Papers, volume 7082 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24869-1 (print), 3-642-24870-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24870-2.

Conati:2007:UMI

[CMP07] Cristina Conati, Kathleen McCoy, and Georgios Paliouras, editors. User Modeling 2007: 11th International Conference, UM 2007, Corfu, Greece, July 25–29, 2007. Proceedings, volume 4511 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73077-X (print), 3-540-73078-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73078-1.

Czumaj:2012:ALPa

[CMPW12a] Artur Czumaj, Kurt Mehlhorn, Andrew Pitts, and Roger Wattenhofer, editors. Automata, Languages, and Programming: 39th International Colloquium, ICALP 2012, Warwick, UK, July 9–13, 2012, Proceedings, Part I, volume 7391 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31593-3 (print), 3-642-31594-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31594-7.

Czumaj:2012:ALPb

[CMPW12b] Artur Czumaj, Kurt Mehlhorn, Andrew Pitts, and Roger Wattenhofer, editors. Automata, Languages, and Programming: 39th International Colloquium, ICALP 2012, Warwick, UK, July 9–13, 2012, Proceedings, Part II, volume 7392 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31584-4 (print), 3-642-31585-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31585-5.

Cunningham:1996:IPC

[CMQ96] William H. Cunningham, S. Thomas McCormick, and Maurice Queyranne, editors. Integer programming and combinatorial optimization: 5th International IPCO Conference, Vancouver, British of Columnia, Canada, June 3–5, 1996: proceedings, volume 1084 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61310-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T57.74 .C62 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1084.htm;
 http://www.springerlink.
com/content/978-3-540-
61310-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1084.

Chen:2008:ISP

[CMS08a] Liqun Chen, Yi Mu, and Willy Susilo, editors. Information Security Practice and Experience: 4th International Conference, IS-PEC 2008 Sydney, Australia, April 21–23, 2008 Proceedings, volume 4991 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79103-5 (print), 3-540-79104-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79104-1.

Cuellar:2008:FFM

[CMS08b] Jorge Cuellar, Tom Maibaum, and Kaisa Sere, editors. FM 2008: Formal Methods: 15th International Symposium on Formal Methods, Turku, Finland, May 26–30, 2008 Proceedings, volume 5014 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68235-X (print), 3-540-68237-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68237-0.

Cohen:2011:ECI

[CMS11] Johanne Cohen, Patrick Maillé, and Burkhard Stiller, editors. Economics of Converged, Internet-Based Networks: 7th International Workshop on Internet Charging and QoS Technologies, ICQT 2011, Paris, France, October 24, 2011. Proceedings, volume 6995 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24546-3 (print), 3-642-24547-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24547-3.

Christianson:2012:SPX

[CMSA12] Bruce Christianson, James Malcolm, Frank Stajano, and Jonathan Anderson, editors. Security Protocols XX: 20th International Workshop, Cambridge, UK, April 12–13, 2012, Revised Selected Papers, volume 7622 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35693-1 (print), 3-642-35694-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35694-0.

Caputo:2010:MCB

[CMSM^{+}10] Barbara Caputo, Henning Müller, Tanveer Syeda-Mahmood, James S.

Duncan, Fei Wang, and Jayashree Kalpathy-Cramer, editors. Medical Content-Based Retrieval for Clinical Decision Support: First MICCAI International Workshop, MCBR-CDS 2009, London, UK, September 20, 2009, Revised Selected Papers, volume 5853 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11768-6 (print), 3-642-11769-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11769-5.

Comon:2001:CCL

[CMT01] Hubert Comon, Claude Marche, and Ralf Treinen, editors. Constraints in computational logics: theory and applications: International Summer School, CCL’99, Gif-sur-Yvette, France, September 5–8, 1999: revised lectures, volume 2002 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41950-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA267.A1 L43 no.2002; QA76.612 .C693 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2002.htm;
 http://www.springerlink.
com/content/978-3-540-
41950-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2002.

Lectures at the summer school organized by the ESPRIT CCL Working Group in Gif-sur-Yvette, France, September 1999.

Carreno:2002:TPH

[CMT02] Victor A. Carreño, Cèsar A.

Muñoz, and Sofiène Tahar, editors. Theorem Proving in Higher Order Logics: 15th International Conference, TPHOLs 2002, Hampton, VA, USA, August 20–23, 2002. Proceedings, volume 2410 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44039-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T655 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2410.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        44039-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2410.

Constantopoulos:1996:AIS

[CMV96] Panos Constantopoulos, John Mylopoulos, and Yannis Vassiliou, editors. Advanced information systems engineering: 8th International Conference CAiSE ’96, Heraklion, Crete, Greece, May 20–24, 1996: proceedings, volume 1080 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61292-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .N67 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1080.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61292-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1080.

Champarnaud:1999:AIT

[CMZ99] Jean-Marc Champarnaud, Denis Maurel, and Djelloul Ziadi, editors. Automata implementation: Third International Workshop on Implementing Automata, WIA’98, Rouen, France, September 17–19, 1998: Revised Papers, volume 1660 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66652-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I5787 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1660.htm;
 http://www.springerlink.
com/content/978-3-540-
66652-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1660.

Chen:2004:ISI

[CMZL04] Hsinchun Chen, Reagan Moore, Daniel D. Zeng, and John Leavitt, editors. Intelligence and Security Informatics: Second Symposium on Intelligence and Security Informatics, ISI 2004, Tucson, AZ, USA, June 10–11, 2004. Proceedings, volume 3073 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22125-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t3073.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22125-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3073;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98042.

Conen:1998:CTC

[CN98] Wolfram Conen and Gustaf Neumann. Coordination technology for collaborative applications: organizations, processes, and agents, volume 1364 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64170-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 282 pp. LCCN QA76.76.D47C667 1998.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1364.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        64170-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1364.

Consens:2005:SPI

[CN05] Mario Consens and Juan A. Navarro González, editors. String processing and information retrieval: 12th international conference, SPIRE 2005, Buenos Aires, Argentina, November 2– 4, 2005: proceedings, volume 3772 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29740-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.T48 I58 2005.

URL http://www.springerlink.
com/content/978-3-540-
29740-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3772.

Christensen:2006:CVS

[CN06] Henrik I. Christensen and Hans-Hellmut Nagel, editors. Cognitive Vision Systems: Sampling the Spectrum of Approaches, volume 3948 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33971-X (print), 3-540-33972-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33972-4.

Chetty:2008:PRB

[CNA08] Madhu Chetty, Alioune Ngom, and Shandar Ahmad, editors. Pattern Recognition in Bioinformatics: Third IAPR International Conference, PRIB 2008, Melbourne, Australia, October 15– 17, 2008. Proceedings, volume 5265 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88434-3 (print), 3-540-88436-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88436-1.

Conte:2005:HPE

[CNmWH^{+}05] Tom Conte, Nacho Navarro, Wen mei W. Hwu, Mateo Valero, and Theo Ungerer, editors. High Performance Embedded Architectures and Compilers: First International Conference, HiPEAC 2005, Barcelona, Spain, November 17–18, 2005. Proceedings, volume 3793 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30317-0 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30317-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3793.

Carpin:2008:SMP

[CNP^{+}08] Stefano Carpin, Itsuki Noda, Enrico Pagello, Monica Reggiani, and Oskar von Stryk, editors. Simulation, Modeling, and Programming for Autonomous Robots: First International Conference, SIMPAR 2008 Venice, Italy, November 3–6, 2008. Proceedings, volume 5325 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89075-0 (print), 3-540-89076-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89076-8.

Cao:2005:APP

[CNX05] Jiannong Cao, Wolfgang (Wolfgang) Nejdl, and Ming Xu, editors. Advanced parallel processing technologies: 6th international workshop, APPT 2005, Hong Kong, China, October 27–28, 2005: proceedings, volume 3756 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29639-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .A67 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29639-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3756.

Ciancarini:1995:OBM

[CNY95] Paolo Ciancarini, Oscar Marius Nierstrasz, and Akinori Yonezawa, editors. Object-based models and languages for concurrent systems: ECOOP ’94 Workshop on Models and Languages for Coordination of Parallelism and Distribution, Bologna, Italy, July 5, 1994: proceedings, volume 924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59450-7 (Berlin), 0-387-59450-7 (New York), 0-387-59450-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E24 1994. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0924.
htm;   http://www.springerlink.
com/content/978-3-540-
59450-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=924.

Carrasco:1994:GIA

[CO94] Rafael C. Carrasco and Jose Oncina, editors. Grammatical inference and applications: second international colloquium, ICGI-94, Alicante, Spain, September 21–23, 1994: proceedings, volume 862 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58473-0 (Berlin), 0-387-58473-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .G67 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0862.htm;
 http://www.springerlink.
com/content/978-3-540-
58473-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=862.

Papers presented during the Second International Colloquium on Grammatical Inference, Alicante, Spain, September 21–23, 1994.

Cohen:2011:SBS

[CÓ11] Myra B. Cohen and Mel Ó Cinnéide, editors. Search Based Software Engineering: Third International Symposium, SSBSE 2011, Szeged, Hungary, September 10– 12, 2011. Proceedings, volume 6956 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23715-0 (print), 3-642-23716-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23716-4.

Cuenca:2006:PWC

[COB06] Pedro Cuenca and Luiz Orozco-Barbosa, editors. Personal Wireless Communications: IFIP TC6 11th International Conference, PWC 2006, Albacete, Spain, September 20–22, 2006. Proceedings, volume 4217 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45174-9 (print), 3-540-45176-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45176-1.

Czarnecki:2008:MDE

[COB^{+}08] Krzysztof Czarnecki, Ileana Ober, Jean-Michel Bruel, Axel Uhl, and Markus Völter, editors. Model Driven Engineering Languages and Systems: 11th International Conference, MoDELS 2008, Toulouse, France, September 28 — October 3, 2008. Proceedings, volume 5301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87874-2 (print), 3-540-87875-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87875-9.

Caivano:2011:PFS

[COBV11] Danilo Caivano, Markku Oivo, Maria Teresa Baldassarre, and Giuseppe Visaggio, editors. Product-Focused Software Process Improvement: 12th International Conference, PROFES 2011, Torre Canne, Italy, June 20– 22, 2011. Proceedings, volume 6759 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21842-3 (print), 3-642-21843-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21843-9.

Codognet:2001:LPI

[Cod01] Philippe Codognet, editor. Logic Programming: 17th International Conference, ICLP 2001, Paphos, Cyprus, November 26–December 1, 2001. Proceedings, volume 2237 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42935-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I56a 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2237.htm;
 http://www.springerlink.
com/content/978-3-540-
42935-7;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2237.

Coelho:1998:PAI

[Coe98] Helder Coelho, editor. Progress in artificial intelligence: Ibero– American Conference on AI, 6th Lisbon, Portugual, October 5– 9, 1998: Proceedings, volume 1484 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64992-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334.I38 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1484.
htm;   http://www.springerlink.
com/content/978-3-540-
64992-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1484.

CoelloCoello:2002:MAA

[Coe02] Carlos A. Coello Coello, editor. MICAI 2002: advances in artificial intelligence: Second Mexican International Conference on Artificial Intelligence, Merida, Yucatan, Mexico, April 22–26, 2002: Proceedings, volume 2313 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43475-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .M48 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2313.
htm;   http://www.springerlink.
com/content/978-3-540-
43475-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2313.

Coello:2011:LIO

[Coe11] Carlos A. Coello Coello, editor. Learning and Intelligent Optimization: 5th International Conference, LION 5, Rome, Italy, January 17–21, 2011. Selected Papers, volume 6683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25565-5 (print), 3-642-25566-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25566-3.

Cohen:1992:ACF

[Coh92] G. Cohen, editor. Algebraic coding: first French-Soviet workshop, Paris, July 22–24, 1991: proceedings, volume 573 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55130-1 (Berlin), 0-387-55130-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A4 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0573.
htm;   http://www.springerlink.
com/content/978-3-540-
55130-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=573.

Cohen:1994:ACF

[Coh94] G. Cohen, editor. Algebraic coding: First French-Israeli Workshop, Paris, France, July 19– 21, 1993: proceedings, volume 781 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-57843-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A397 1994. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0781.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57843-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=781.

Cohen:1996:ANT

[Coh96] Henri Cohen, editor. Algorithmic number theory: second international symposium, ANTS-II, Talence, France, May 18–23, 1996: proceedings, volume 1122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61581-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA241 .A43 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1122.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61581-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1122.

Cohen:2010:PPC

[Coh10] David Cohen, editor. Principles and Practice of Constraint Programming – CP 2010: 16th International Conference, CP 2010, St. Andrews, Scotland, September 6–10, 2010. Proceedings, volume 6308 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15395-X (print), 3-642-15396-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15396-9.

Cointe:1996:EOO

[Coi96] Pierre Cointe, editor.

ECOOP ’96 — objectoriented programming: 10th European Conference, Linz, Austria, July 8–12, 1996: proceedings, volume 1098 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61439-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1098.htm;
 http://www.springerlink.
com/content/978-3-540-
61439-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1098.

Cointe:1999:MLA

[Coi99] Pierre Cointe, editor. Meta-level architectures and reflection: Second International Conference, Reflection’99, Saint-Malo, France, July 19–21, 1999: Proceedings, volume 1616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66280-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .R46 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1616.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66280-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1616.

Coleman:1984:LSN

[Col84] Thomas F. (Thomas Frederick) Coleman. Large sparse numerical optimization, volume 165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-12914-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). iv + 105 pp. LCCN QA402.5 .C543 1984. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0165.htm;
http://www.springerlink.
com/content/978-0-387-
12914-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=165.

Collet:2002:AEI

[Col02] Pierre Collet, editor. Artificial evolution: 5th International Conference, Evolution Artificielle, EA 2001, Le Creusot, France, October 29–31, 2001: Selected Papers, volume 2310 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43544-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .A217 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2310.htm;
http://www.springerlink.
com/content/978-3-540-
43544-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2310.

Collard:2007:OBD

[Col07] Martine Collard, editor. Ontologies-Based Databases and Information Systems: First and Second VLDB Workshops, ODBIS 2005/ 2006 Trondheim, Norway, September 2–3, 2005 Seoul, Korea, September 11, 2006 Revised Papers, volume 4623 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75473-3 (print), 3-540-75474-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75474-9.

Comon:1997:RTA

[Com97] Hubert Comon, editor. Rewriting techniques and applications: 8th International Conference, RTA-97, Sitges, Spain, June 2–5, 1997: proceedings, volume 1232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62950-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .R48 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1232.htm;

http://www.springerlink.
com/content/978-3-540-
62950-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1232.

Carrasco-Ochoa:2012:PRM

[COMTLB12] Jesús Ariel Carrasco-Ochoa, José Francisco Martínez-Trinidad, José Arturo Olvera López, and Kim L. Boyer, editors. Pattern Recognition: 4th Mexican Conference, MCPR 2012, Huatulco, Mexico, June 27–30, 2012. Proceedings, volume 7329 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31148-2 (print), 3-642-31149-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31149-9.

Conradi:1997:SCM

[Con97] Reidar Conradi, editor. Software configuration management: ICSE ’97 SCM-7 Workshop, Boston, MA, USA, May 18–19, 1997: proceedings, volume 1235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63014-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C69 I25 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1235.htm;
 http://www.springerlink.
com/content/978-3-540-
63014-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1235.

Conradi:2000:SPT

[Con00] Reidar Conradi, editor. Software process technology: 7th European workshop, EWSPT 2000, Kaprun, Austria, February 21–25, 2000: Proceedings, volume 1780 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67140-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1780.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1780.
htm;   http://www.springerlink.
com/content/978-3-540-
67140-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1780.

Coppersmith:1995:ACC

[Cop95] Don Coppersmith, editor. Advances in cryptology, CRYPTO ’95: 15th Annual International Cryptology Conference, Santa Barbara, California, USA, August 27–31, 1995: proceedings, volume 963 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60221-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0963.htm;
http://www.springerlink.
com/content/978-3-540-
60221-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=963.

“Sponsored by the International Association for Cryptologic Research (IACR), in cooperation with the IEEE Computer Society Technical Committee on Security and Privacy”–Pref.

Cortesi:2002:VMC

[Cor02] Agostino Cortesi, editor. Verification, model checking and abstract interpretation: Third International Workshop, VMCAI 2002, Venice, Italy, January 21–22, 2002: Revised Papers, volume 2294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43631-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 V53 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2294.htm;
 http://www.springerlink.
com/content/978-3-540-
43631-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2294.

Caromel:1998:COO

[COT98] Denis Caromel, Rodney R. Oldehoeft, and Marydell Tholburn, editors. Computing in object-oriented parallel environments: Second International Symposium, ISCOPE 98, Santa Fe, NM, USA, December 8–11, 1998: Proceedings, volume 1505 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65387-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .I83 1998; QA267.A1 L43 no.1505. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1505.htm;
 http://www.springerlink.
com/content/978-3-540-
65387-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1505.

Courcoubetis:1993:CAV

[Cou93a] Costas Courcoubetis, editor. Computer aided verification: 5th international conference, CAV’93, Elounda, Greece, June 28–July 1, 1993: proceedings, volume 697 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56922-7 (Berlin), 0-387-56922-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E93 C38 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0697.htm;
 http://www.springerlink.
com/content/978-3-540-
56922-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=697.

Cousot:1993:SAT

[Cou93b] Patrick Cousot, editor. Static analysis: third International Workshop, WSA ’93, Padova, Italy, September 1993: proceedings, volume 724 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57264-3 (USA), 3-540-57264-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .W6985 1993. DM58.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0724.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57264-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=724.

Cousot:2001:SAI

[Cou01] Patrick Cousot, editor. Static analysis: 8th International Symposium, SAS 2001,Paris, France, July 16–18, 2001: Proceedings, volume 2126 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42314-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.6 .I5733 2001; QA267.A1 L43 no.2126. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2126.htm;
http://www.springerlink.
com/content/978-3-540-
42314-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2126.

Cousot:2003:SAI

[Cou03] Radhia Cousot, editor. Static analysis: 10th International Symposium, SAS 2003, San Diego, CA, USA, June 11–13, 2003: Proceedings, volume 2694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40325-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2694.htm;
http://www.springerlink.
com/content/978-3-540-
40325-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2694.

Cousot:2005:VMC

[Cou05] Radhia Cousot, editor. Verification, Model Checking, and Abstract Interpretation: 6th International Conference, VMCAI 2005, Paris, France, January 17– 19, 2005. Proceedings, volume 3385 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24297-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 V53.

URL http://www.springerlink.
com/content/978-3-540-
24297-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3385;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105073.

Covaci:1999:ANF

[Cov99] Stefan Covaci, editor. Active networks: First International Workshop, IWAN’99, Berlin, Germany, June 30–July 2, 1999: Proceedings, volume 1653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66238-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I95 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1653.htm;
http://www.springerlink.
com/content/978-3-540-
66238-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1653.

Cowell:1977:PNS

[Cow77] Wayne R. Cowell, editor.

Portability of numerical software workshop, Oak Brook, Illinois, June 21– 23, 1976, volume 57 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CODEN LNCSD9. ISBN 0-387-08446-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA297 .W65 1976. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0057.htm;
http://www.springerlink.
com/content/978-0-387-
08446-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=57.

Workshop organized by Argonne National Laboratory.

Chaum:1988:ACE

[CP88] David Chaum and Wyn L.

Price, editors. Advances in cryptology — EU-ROCRYPT ’87: Workshop on the Theory and Application of Cryptographic Techniques, Amsterdam, The Netherlands, April 13–15, 1987: proceedings, volume 304 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19102-X (New York), 3-540-19102-X (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E9631 1987; QA267.A1 L43 no.304. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0304.htm;
 http://www.springerlink.
com/content/978-0-387-
19102-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=304.

Sponsored by the International Association for Crytologic Research.

Calmet:1998:AIS

[CP98a] Jacques Calmet and Jan Plaza, editors. Artificial intelligence and symbolic computation: International Conference, AISC’98, Plattsburgh, New York, USA, September 16–18, 1998: Proceedings, volume 1476 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64960-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1476. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1476.htm;
 http://www.springerlink.
com/content/978-3-540-
64960-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1476.

Chin:1998:CVAa

[CP98b] Roland T. Chin and Ting-Chuen Pong, editors. Computer vision, ACCV’98: Third Asian Conference on Computer Vision, Hong Kong, China, January 8– 10, 1998: Proceedings. Volume 1, volume 1351 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 0-387-63930-6, 3-540-63930-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634.A85 1998; QA267.A1 L43 no.1351-1352. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1351.
htm;   http://www.springerlink.

                                        com/content/978-0-387-
                                        63930-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1351.

Chin:1998:CVAb

[CP98c] Roland T. Chin and Ting-Chuen Pong, editors. Computer vision, ACCV’98: Third Asian Conference on Computer Vision, Hong Kong, China, January 8– 10, 1998: Proceedings. Volume 2, volume 1352 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-63931-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634.A85 1998; QA267.A1 L43 no.1351-1352. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1352.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63931-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1352.

Ciobanu:1999:FCT

[CP99] Gabriel Ciobanu and Gheorghe Păun, editors. Fundamentals of computation theory: 12th International Symposium, FCT’99, Iasi, Romania, August 30– September 3, 1999: Proceedings, volume 1684 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66412-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1684; QA75.5 .F367 1999. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1684.
htm;   http://www.springerlink.
com/content/978-3-540-
66412-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1684.

Craw:2002:ACB

[CP02] Susan Craw and Alun Preece, editors. Advanced in Case-Based Reasoning: 6th European Conference, EC-CBR 2002, Aberdeen, Scotland, UK, September 4–7, 2002. Proceedings, volume 2416 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44109-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E935 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2416.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44109-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2416.

Costabile:2005:HCI

[CP05] Maria Francesca Costabile and Fabio Paternò, editors. Humancomputer interaction: INTERACT 2005: IFIP TC13 international conference, Rome, Italy, September 12–16, 2005: proceedings, volume 3585 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28943-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I584 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        28943-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3585.

Carbone:2006:DCI

[CP06a] Alessandra Carbone and Niles A. Pierce, editors. DNA computing: 11th international workshop on DNA computing, DNA11, London, on, Canada, June 6– 9, 2005. revised selected papers, volume 3892 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-34161-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34161-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3892.

Cheng:2006:AAI

[CP06b] Siu-Wing Cheng and Chung Keung Poon, editors. Algorithmic Aspects in Information and Management: Second International Conference, AAIM 2006, Hong Kong, China, June 20–22, 2006. Proceedings, volume 4041 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35157-4 (print), 3-540-35158-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35158-0.

Cook:2007:VMC

[CP07] Byron Cook and Andreas Podelski, editors. Verification, Model Checking, and Abstract Interpretation: 8th International Conference, VM-CAI 2007, Nice, France, January 14–16, 2007. Proceedings, volume 4349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69735-7 (print), 3-540-69738-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69738-1.

Carlet:2010:STA

[CP10a] Claude Carlet and Alexander Pott, editors. Sequences and Their Applications — SETA 2010: 6th International Conference, Paris, France, September 13–17, 2010. Proceedings, volume 6338 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15873-0 (print), 3-642-15874-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15874-2.

Carro:2010:PAD

[CP10b] Manuel Carro and Ricardo Peña, editors. Practical Aspects of Declarative Languages: 12th International Symposium, PADL 2010, Madrid, Spain, January 18–19, 2010. Proceedings, volume 5937 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11502-0 (print), 3-642-11503-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11503-5.

Cimiano:2010:KEM

[CP10c] Philipp Cimiano and H. Sofia Pinto, editors. Knowledge Engineering and Management by the Masses: 17th International Conference, EKAW 2010, Lisbon, Portugal, October 11–15, 2010. Proceedings, volume 6317 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16437-4 (print), 3-642-16438-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16438-5.

Cerone:2011:TAC

[CP11] Antonio Cerone and Pekka Pihlajasaari, editors. Theoretical Aspects of Computing — ICTAC 2011: 8th International Colloquium, Johannesburg, South Africa, August 31 — September 2, 2011. Proceedings, volume 6916 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23282-5 (print), 3-642-23283-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23283-1.

Carbone:2012:WSF

[CP12] Marco Carbone and Jean-Marc Petit, editors. Web Services and Formal Methods: 8th International Workshop, WS-FM 2011, Clermont-Ferrand, France, September 1–2, 2011, Revised Selected Papers, volume 7176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29833-8 (print), 3-642-29834-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29834-9.

Chatterjee:1999:LCP

[CPC^{+}99] Siddhartha Chatterjee, Jan F. Prins, Larry Carter, Jeanne Ferrante, Zhiyuan Li, David Sehr, and Pen-Chung Yew, editors. Languages and compilers for parallel computing: 11th International Workshop, LCPC ’98, Chapel Hill, NC, USA August 7–9, 1998: Proceedings, volume 1656 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66426-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1656; QA76.58 .W656 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1656.
htm;   http://www.springerlink.
com/content/978-3-540-
66426-0;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1656.

Choi:2007:ACS

[CPC07] Lynn Choi, Yunheung Paek, and Sangyeun Cho, editors. Advances in Computer Systems Architecture: 12th Asia-Pacific Conference, AC-SAC 2007, Seoul, Korea, August 23–25, 2007. Proceedings, volume 4697 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74308-1 (print), 3-540-74309-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74309-5.

Cantu-Paz:2003:GECa

[CPFD^{+}03a] Erick Cantú-Paz, James A. Foster, Kalyanmoy Deb, Lawrence David Davis, Rajkumar Roy, Una-May O’Reilly, Hans-Georg Beyer, Russell Standish, Graham Kendall, Stewart Wilson, Mark Harman, Joachim Wegener, Dipankar Dasgupta, Mitch A. Potter, Alan C. Schultz, Kathryn A. Dowsland, Natasha Jonoska, and Julian Miller, editors.

Genetic and Evolutionary Computation GECCO 2003 Genetic and Evolutionary Computation Conference Chicago, IL, USA, July 12–16, 2003 Proceedings, Part I, volume 2723. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40602-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA402.5 .G4563 2003. URL http:

//www.springerlink.
com/content/978-3-540-
40602-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2723.

Cantu-Paz:2003:GECb

[CPFD^{+}03b] Erick Cantú-Paz, James A.

Foster, Kalyanmoy Deb, Lawrence David Davis, Rajkumar Roy, Una-May O’Reilly, Hans-Georg Beyer, Russell Standish, Graham Kendall, Stewart Wilson, Mark Harman, Joachim Wegener, Dipankar Dasgupta, Mitch A. Potter, Alan C. Schultz, Kathryn A. Dowsland, Natasha Jonoska, and Julian Miller, editors. Genetic and Evolutionary Computation GECCO 2003 Genetic and Evolutionary Computation Conference Chicago, IL, USA, July 12–16, 2003 Proceedings, Part II, volume 2724. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40603-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA402.5 .G4563 2003. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        40603-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2724.

Chen:2005:PDP

[CPGL05] Guihai Chen, Yi Pan, Minyi Guo, and Jian Lu, editors. Parallel and distributed processing and applications: ISPA 2005 workshops: ISPA 2005 international workshops AEPP, ASTD, BIOS, GCIC, IADS, MASN, SGCA, and WISA, Nanjing, China, November 2–5, 2005: proceedings, volume 3759 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29770-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P367 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29770-3;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3759.

Crescenzi:2007:FAI

[CPP07] Pierluigi Crescenzi, Giuseppe Prencipe, and Geppino Pucci, editors. Fun with Algorithms: 4th International Conference, FUN 2007, Castiglioncello, Italy, June 3–5, 2007. Proceedings, volume 4475 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72913-5 (print), 3-540-72914-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72914-3.

Clark:2006:SPC

[CPPB06] John A. Clark, Richard F.

Paige, Fiona A. C. Polack, and Phillip J. Brooke, editors. Security in Pervasive Computing: Third International Conference, SPC 2006, York, UK, April 18–21, 2006. Proceedings, volume 3934 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33376-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33376-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3934.

Camara:2010:SAC

[CPR^{+}10] Oscar Camara, Mihaela Pop, Kawal Rhode, Maxime Sermesant, Nic Smith, and Alistair Young, editors. Statistical Atlases and Computational Models of the Heart: First International Workshop, STA-COM 2010, and Cardiac Electrophysiological Simulation Challenge, CESC 2010, Held in Conjunction with MICCAI 2010, Beijing, China, September 20, 2010. Proceedings, volume 6364 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15834-X (print), 3-642-15835-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15835-3.

Cox:2012:DRI

[CPR12] Philip Cox, Beryl Plimmer, and Peter Rodgers, editors. Diagrammatic Representation and Inference: 7th International Conference, Diagrams 2012, Canterbury, UK, July 2–6, 2012. Proceedings, volume 7352 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31222-5 (print), 3-642-31223-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31223-6.

Calude:2001:MPM

[CPRS01] Cristian S. Calude, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa, editors. Multiset Processing: Mathematical, Computer Science, and Molecular Computing Points of View, volume 2235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43063-6. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 358 pp. LCCN QA76.887 .M85 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2235.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        43063-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2235.

Cabestany:2005:CIB

[CPS05] Joan Cabestany, Alberto Prieto, and Francisco Sandoval, editors. Computational Intelligence and Bioinspired Systems: 8th International Workshop on Artificial Neural Networks, IWANN 2005, Vilanova i la Geltrú, Barcelona, Spain, June 8–10, 2005. Proceedings, volume 3512 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26208-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I5845 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26208-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3512.

Casanovas:2010:AAC

[CPSA10] Pompeu Casanovas, Ugo Pagallo, Giovanni Sartor, and Gianmaria Ajani, editors. AI Approaches to the Complexity of Legal Systems. Complex Systems, the Semantic Web, Ontologies, Argumentation, and Dialogue: International Workshops AICOL-I/ IVR-XXIV Beijing, China, September19, 2009 and AICOL-II/JURIX 2009, Rotterdam,The Netherlands, December 16, 2009 Revised Selected Papers, volume 6237 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16523-0 (print), 3-642-16524-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16524-5.

Cechich:2003:CBS

[CPV03] Alejandra Cechich, Mario Piattini, and Antonio Vallecillo, editors. Component-Based Software Quality: Methods and Techniques, volume 2693 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40503-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 402 pp. LCCN QA76.76.C66 C66 2003.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t2693.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40503-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2693;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11721.

Coen-Porisini:2003:SEM

[CPvdH03] Alberto Coen-Porisini and Andre van der Hoek, editors. Software engineering and middleware: Third International Workshop, SEM 2002, Orlando, FL, USA, May 20–21, 2002: Revised Papers, volume 2596 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-07549-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S455 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2596.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        07549-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2596.

Chen:2007:CAP

[CPZ07] Bo Chen, Mike Paterson, and Guochuan Zhang, editors. Combinatorics, Algorithms, Probabilistic and Experimental Methodologies: First International Symposium, ESCAPE 2007, Hangzhou, China, April 7–9, 2007, Revised Selected Papers, volume 4614 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74449-5 (print), 3-540-74450-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74450-4.

Choudhury:2009:LCA

[CQSS09] Tanzeem Choudhury, Aaron Quigley, Thomas Strang, and Koji Suginuma, editors. Location and Context Awareness: 4th International Symposium, LoCA 2009 Tokyo, Japan, May 7–8, 2009 Proceedings, volume 5561 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01720-7 (print), 3-642-01721-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01721-6.

Conti:1973:COTa

[CR73a] Roberto Conti and Antonio Ruberti, editors. 5th Conference on Optimization Techniques, Rome, Italy, May 7–11, 1973. Part I, volume 3 of Series, I.F.I.P. TC 7 optimization conferences and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1973. CO-DEN LNCSD9. ISBN 0-387-06583-0 (New York), 3-540-06583-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.3–4. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0003.htm;
 http://www.springerlink.
com/content/978-0-387-
06583-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3. En-

glish or French. Held May 7–11, 1973; sponsored by the IFIP Technical Committee on Optimization (TC-7) and the Italian National Research Council.

Conti:1973:COTb

[CR73b] Roberto Conti and Antonio Ruberti, editors. 5th Conference on Optimization Techniques, Rome, Italy, May 7–11, 1973. Part II, volume 4 of Series, I.F.I.P. TC 7 optimization conferences and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1973. CODEN LNCSD9. ISBN 3-540-06600-4, 0-387-06600-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.3–4. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0004.htm;
http://www.springerlink.
com/content/978-3-540-
06600-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=4. En-

glish or French. Held May 7–11, 1973; sponsored by the IFIP Technical Committee on Optimization (TC-7) and the Italian National Research Council.

Cerioli:2001:RTA

[CR01a] Maura Cerioli and Gianna Reggio, editors. Recent Trends in Algebraic Development Techniques: 15th International Workshop, WADT 2001, Joint with the CoFI WG Meeting, Genova, Italy, April 1–3, 2001. Selected Papers, volume 2267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-43159-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 W33 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2267.
htm;   http://www.springerlink.
com/content/978-3-540-
43159-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2267.

Condon:2001:DCI

[CR01b] Anne Condon and Grzegorz Rozenberg, editors. DNA computing: 6th International Workshop on DNA-Based Computers, DNA 2000, Leiden, the Netherlands, June 13– 17, 2000: Revised Papers, volume 2054 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42076-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2054; QA76.887 .I58 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2054.htm;
 http://www.springerlink.
com/content/978-3-540-
42076-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2054.

Chen:2004:DCI

[CR04a] Junghuei Chen and John Reif, editors. DNA Computing: 9th International Workshop on DNA Based Computers, DNA9, Madison, WI, USA, June 1–3, 2003: Revised Papers, volume 2943 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20930-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.887 .I58 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2943.htm;
 http://www.springerlink.
com/content/978-3-540-
20930-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2943;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95518.

Cortadella:2004:ATP

[CR04b] Jordi Cortadella and Wolfgang Reisig, editors. Applications and Theory of Petri Nets 2004: 25th International Conference, ICATPN 2004, Bologna, Italy, June 21–25, 2004. Proceedings, volume 3099 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22236-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3099.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22236-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3099;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98283.

Castro:2005:PPS

[CR05a] Miguel Castro and Robbert Renesse, editors. Peer-to-peer systems IV: 4th international workshop, IPTPS 2005, Ithaca, NY, USA, February 24– 25, 2005. revised selected papers, volume 3640 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29068-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29068-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3640.

Crestani:2005:ICN

[CR05b] Fabio Crestani and Ian Ruthven, editors. Information Context: Nature, Impact, and Role: 5th International Conference on Conceptions of Library and Information Sciences, Co-LIS 2005 Glasgow, UK, June 4–8, 2005 Proceedings, volume 3507 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26178-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26178-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3507;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b137025.

Cramer:2005:ACE

[Cra05] Ronald Cramer, editor. Advances in cryptology: EUROCRYPT 2005: 24th Annual International Conference on the Theory and Applications of Cryptographic Techniques, Aarhus, Denmark, May 22–26, 2005, proceedings, volume 3494 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25910-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25910-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3494;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136415.

Cramer:2008:PKC

[Cra08] Ronald Cramer, editor. Public Key Cryptography — PKC 2008: 11th International Workshop on Practice and Theory in Public-Key Cryptography, Barcelona, Spain, March 9–12, 2008. Proceedings, volume 4939 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78439-X (print), 3-540-78440-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78440-1.

Cramer:2012:TCT

[Cra12] Ronald Cramer, editor.

Theory of Cryptography: 9th Theory of Cryptography Conference, TCC 2012, Taormina, Sicily, Italy, March 19–21. Proceedings, volume 7194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28913-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28913-2.

Chowdhury:2008:PCI

[CRD08] Dipanwita Roy Chowdhury, Vincent Rijmen, and Abhijit Das, editors. Progress in Cryptology — INDOCRYPT 2008: 9th International Conference on Cryptology in India, Kharagpur, India, December 14–17, 2008. Proceedings, volume 5365 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89753-4 (print), 3-540-89754-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89754-5.

Cabestany:2011:ACIa

[CRJ11a] Joan Cabestany, Ignacio Rojas, and Gonzalo Joya, editors. Advances in Computational Intelligence: 11th International Work-Conference on Artificial Neural Networks, IWANN 2011, Torremolinos-Málaga, Spain, June 8–10, 2011, Proceedings, Part I, volume 6691 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21500-9 (print), 3-642-21501-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21501-8.

Cabestany:2011:ACIb

[CRJ11b] Joan Cabestany, Ignacio Rojas, and Gonzalo Joya, editors. Advances in Computational Intelligence: 11th International Work-Conference on Artificial Neural Networks, IWANN 2011, Torremolinos-Málaga, Spain, June 8–10, 2011, Proceedings, Part II, volume 6692 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21497-5 (print), 3-642-21498-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21498-1.

Campbell:2001:AIS

[CRL01] J. A. Campbell and Eugenio Roanes-Lozano, editors. Artificial intelligence and symbolic computation: International Conference AISC 2000, Madrid, Spain, July 17– 19, 2000: Revised Papers, volume 1930 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42071-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I5488 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1930.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42071-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1930.

Crowley:2003:CVS

[Cro03] J. L. Crowley, editor. Computer vision systems: Third International Conference, ICVS 2003, Graz, Austria, April 1–3, 2003: Proceedings, volume 2626 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00921-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .I27 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2626.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00921-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2626.

Crowcroft:2000:QFI

[CRS00] Jon Crowcroft, James Roberts, and Mikhail I. Smirnov, editors. Quality of Future Internet Services: First COST 263 International Workshop, QofIS 2000, Berlin, Germany, September 25– 26, 2000. Proceedings, volume 1922 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41076-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 Q36 2000; QA267.A1 L43 no.1922. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1922.htm;
http://www.springerlink.
com/content/978-3-540-
41076-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1922.

Corchado:2010:HAI

[CRS10] Emilio Corchado, Manuel Graña Romay, and Alexandre Manhaes Savio, editors. Hybrid Artificial Intelligence Systems: 5th International Conference, HAIS 2010, San Sebastián, Spain, June 23–25, 2010. Proceedings, Part II, volume 6077 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13802-0 (print), 3-642-13803-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13803-4.

Calude:2011:RCS

[CRS11] Cristian S. Calude, Grzegorz Rozenberg, and Arto Salomaa, editors. Rainbow of Computer Science: Dedicated to Hermann Maurer on the Occasion of His 70th Birthday, volume 6570 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19390-0 (print), 3-642-19391-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19391-0.

Cavedon:1997:IAS

[CRW97] Lawrence Cavedon, Anand Rao, and Wayne Wobcke, editors. Intelligent agent systems: theoretical and practical issues: based on a workshop held at PRI-CAI’96, Cairns, Australia, August 26–30, 1996, volume 1209 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62686-7 (Berlin softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 I55 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1209.
htm;   http://www.springerlink.
com/content/978-3-540-
62686-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1209.

Chen:2008:ICS

[CRW08] Liqun Chen, Mark D.

Ryan, and Guilin Wang, editors. Information and Communications Security: 10th International Conference, ICICS 2008 Birmingham, UK, October 20 — 22, 2008 Proceedings, volume 5308 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88624-9 (print), 3-540-88625-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88625-9.

Croitoru:2012:GSK

[CRW^{+}12] Madalina Croitoru, Sebastian Rudolph, Nic Wilson, John Howse, and Olivier Corby, editors. Graph Structures for Knowledge Representation and Reasoning: Second International Workshop, GKR 2011, Barcelona, Spain, July 16, 2011. Revised Selected Papers, volume 7205 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29448-0 (print), 3-642-29449-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29449-5.

Cremers:2009:SGA

[CRYS09] Daniel Cremers, Bodo Rosenhahn, Alan L. Yuille, and Frank R. Schmidt. Statistical and geometrical approaches to visual motion analysis: International Dagstuhl Seminar, Dagstuhl Castle, July 13– 18, 2008: revised papers, volume 5604 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-642-03061-0 (e-book), 3-642-03060-2. viii + 322 pp. LCCN TA1634 .I62 2008. URL http:

//www.springerlink.
com/content/978-3-642-
03061-1.

Cesarini:1987:DMP

[CS87] Francesca Cesarini and Silvio Salza. Database machine performance: modeling methodologies and evaluation strategies, volume 257 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-17942-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 250 pp. LCCN QA76.9.D3 D358561 1987.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0257.
htm;   http://www.springerlink.
com/content/978-0-387-
17942-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=257.

Cousot:1996:SAT

[CS96] Radhia Cousot and David A. Schmidt, editors. Static analysis: Third International Symposium, SAS ’96, Aachen, Germany, September 24–26, 1996: proceedings, volume 1145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61739-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1145.htm;
 http://www.springerlink.
com/content/978-3-540-
61739-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1145.

Corne:1997:ECA

[CS97] David Corne and Jonathan L. Shapiro, editors. Evolutionary computing: AISB International Workshop, Manchester, UK, April 7–8, 1997: proceedings, volume 1305 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63476-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618.A22 1997. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t1305.htm;
http://www.springerlink.
com/content/978-3-540-
63476-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1305.

Caferra:2000:ADC

[CS00a] Ricardo Caferra and Gernot Salzer, editors. Automated Deduction in Classical and Non-Classical Logics: Selected Papers, volume 1761 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67190-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 A85 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1761.htm;
http://www.springerlink.
com/content/978-3-540-
67190-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1761.

Clote:2000:CSL

[CS00b] Peter Clote and Helmut Schwichtenberg, editors. Computer science logic: 14th International Workshop, CSL 2000, Annual Conference of the EACSL, Fischbachau, Germany, August 21–26, 2000: Proceedings, volume 1862 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67895-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.L63 W67 2000; QA75.5 .W624 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1862.
htm;   http://www.springerlink.
com/content/978-3-540-
67895-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1862.

Constantopoulos:2001:RAT

[CS01a] Panos Constantopoulos and Ingeborg T. Sølvberg, editors. Research and advanced technology for digital libraries: 5th European Conference, ECDL 2001, Darmstadt, Germany, September 4–9, 2001: Proceedings, volume 2163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42537-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z692.C65 E28 2001; QA267.A1 L43 no.2163. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2163.htm;
 http://www.springerlink.
com/content/978-3-540-
42537-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2163.

Craeynest:2001:RST

[CS01b] Dirk Craeynest and Alfred Strohmeier, editors. Reliable software technologies: Ada-Europe 2001: 6th Ada-Europe International Conference on Reliable Software Technologies, Leuven, Belgium, May 14– 18, 2001: Proceedings, volume 2043 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42123-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 2001; QA267.A1 L43 no.2043. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2043.htm;
 http://www.springerlink.
com/content/978-3-540-
42123-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2043.

Cohen:2002:AAI

[CS02a] Robin Cohen and Bruce Spencer, editors. Advances in artificial intelligence: 15th Conference of the Canadian Society for Computational Studies of Intelligence, AI 2002, Calgary, Canada, May 27–29, 2002: Proceedings, volume 2338 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43724-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .C36 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2338.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43724-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2338.

Cook:2002:IPC

[CS02b] William J. Cook and Andreas S. Schulz, editors. Integer Programming and Combinatorial Optimization: 9th International IPCO Conference, Cambridge, MA, USA, May 27–29, 2002. Proceedings, volume 2337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43676-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T57.74 .C62 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2337.htm;
 http://www.springerlink.
com/content/978-3-540-
43676-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2337.

Carbonell:2003:UAT

[CS03] Noelle Carbonell and Constantine Stephanidis, editors. Universal access: theoretical perspectives, practice, and experience: 7th ERCIM International Workshop on User Interfaces for All, Paris, France, October 24–25, 2002: Revised Papers, volume 2615 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00855-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.U83 E73 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2615.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00855-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2615.

Cardoso:2005:SWS

[CS05a] Jorge Cardoso and Amit Sheth, editors. Semantic Web Services and Web Process Composition: First International Workshop, SWSWPC 2004, San Diego, CA, USA, July 6, 2004. Revised Selected Papers, volume 3387 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24328-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815 .S98 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24328-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3387;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105145.

Christensen:2005:IPM

[CS05b] Gary E. Christensen and Milan Sonka, editors. Information Processing in Medical Imaging: 19th International Conference, IPMI 2005, Glenwood Springs, CO, USA, July 10–15, 2005. Proceedings, volume 3565 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26545-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN RC78.7.D53 I57 2005.

URL http://www.springerlink.
com/content/978-3-540-
26545-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3565.

Carlet:2007:AFF

[CS07] Claude Carlet and Berk Sunar, editors. Arithmetic of Finite Fields: First International Workshop, WAIFI 2007, Madrid, Spain, June 21– 22, 2007. Proceedings, volume 4547 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73073-7 (print), 3-540-73074-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73074-3.

Choppy:2010:FCS

[CS10] Christine Choppy and Oleg Sokolsky, editors. Foundations of Computer Software. Future Trends and Techniques for Development: 15th Monterey Workshop 2008, Budapest, Hungary, September 24–26, 2008, Revised Selected Papers, volume 6028 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12565-4 (print), 3-642-12566-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12566-9.

Cardelli:2011:DCM

[CS11] Luca Cardelli and William Shih, editors. DNA computing and molecular programming: 17 International Conference, DNA 17, Pasadena, CA, USA, September 19– 23, 2011. proceedings, volume 6937 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23637-5 (soft cover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23637-2;   http://www.
springerlink.com/content/
978-3-642-23637-2/.

Cimatti:2012:TAS

[CS12a] Alessandro Cimatti and Roberto Sebastiani, editors. Theory and Applications of Satisfiability Testing – SAT 2012: 15th International Conference, Trento, Italy, June 17–20, 2012. Proceedings, volume 7317 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31611-5 (print), 3-642-31612-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31612-8.

Constable:2012:LPS

[CS12b] Robert L. Constable and Alexandra Silva, editors. Logic and Program Semantics: Essays Dedicated to Dexter Kozen on the Occasion of His 60th Birthday, volume 7230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29484-7 (print), 3-642-29485-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29485-3.

Cranefield:2012:ABS

[CS12c] Stephen Cranefield and Insu Song, editors. Agent Based Simulation for a Sustainable Society and Multi-agent Smart Computing: International Workshops, PRIMA 2011, Wollongong, Australia, November 14, 2011 Revised Selected Papers, volume 7580 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35611-7 (print), 3-642-35612-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        35612-4.

Corchado:2012:HAIa

[CSA^{+}12a] Emilio Corchado, Václav Snášel, Ajith Abraham, Michał Woźniak, Manuel Graña, and Sung-Bae Cho, editors. Hybrid Artificial Intelligent Systems: 7th International Conference, HAIS 2012, Salamanca, Spain, March 28–30th, 2012. Proceedings, Part I, volume 7208 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28941-X (print), 3-642-28942-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28942-2.

Corchado:2012:HAIb

[CSA^{+}12b] Emilio Corchado, Václav Snášel, Ajith Abraham, Michał Woźniak, Manuel Graña, and Sung-Bae Cho, editors. Hybrid Artificial Intelligent Systems: 7th International Conference, HAIS 2012, Salamanca, Spain, March 28–30th, 2012. Proceedings, Part II, volume 7209 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28930-4 (print), 3-642-28931-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        28931-6.

Combi:2009:AIM

[CSAH09] Carlo Combi, Yuval Shahar, and Ameen Abu-Hanna, editors. Artificial Intelligence in Medicine: 12th Conference on Artificial Intelligence in Medicine, AIME 2009, Verona, Italy, July 18– 22, 2009. Proceedings, volume 5651 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02975-2 (print), 3-642-02976-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02976-9.

Cairo:2000:MAA

[CSCO00] Osvaldo Cairo, L. Enrique Sucar, and Francisco J. Cantu-Ortiz, editors. MICAI 2000: Advances in artificial intelligence: Mexican International Conference on Artificial Intelligence, Acapulco, Mexico, April 2000: Proceedings, volume 1793 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67354-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1793.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1793.
htm;   http://www.springerlink.
com/content/978-3-540-
67354-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1793.

Casanovas:2008:CML

[CSCR08] Pompeu Casanovas, Giovanni Sartor, Núria Casellas, and Rossella Rubino, editors. Computable Models of the Law: Languages, Dialogues, Games, Ontologies, volume 4884 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85568-8 (print), 3-540-85569-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85569-9.

Cao:2006:MAH

[CSJD06] Jiannong Cao, Ivan Stojmenovic, Xiaohua Jia, and Sajal K. Das, editors. Mobile Ad-hoc and Sensor Networks: Second International Conference, MSN 2006, Hong Kong, China, December 13–15, 2006. Proceedings, volume 4325 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49932-6 (print), 3-540-49933-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49933-6.

Chai:2010:ASC

[CSM10] Sun-Ki Chai, John J. Salerno, and Patricia L. Mabry, editors. Advances in Social Computing: Third International Conference on Social Computing, Behavioral Modeling, and Prediction, SBP 2010, Bethesda, MD, USA, March 30–31, 2010. Proceedings, volume 6007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12078-4 (print), 3-642-12079-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12079-4.

Christin:2014:FCD

[CSN14] Nicolas Christin and Reihaneh Safavi-Naini, editors. Financial Cryptography and Data Security: FC 2014 Workshops, BITCOIN and WAHC 2014, Christ Church, Barbados, March 7, 2014, Revised Selected Papers, volume 8438 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2014. ISBN 3-662-44773-8 (print), 3-662-44774-6 (e-book). LCCN QA76.9.A25. URL

http://nbn-resolving.
de/urn:nbn:de:1111-
20141008380;   http:
//d-nb.info/1059163349/
34;   http://www.springerlink.
com/content/978-3-662-
44774-1.

Cabestany:2009:BIS

[CSPC09] Joan Cabestany, Francisco Sandoval, Alberto Prieto, and Juan M. Corchado, editors. Bio-Inspired Systems: Computational and Ambient Intelligence: 10th International Work-Conference on Artificial Neural Networks, IWANN 2009, Salamanca, Spain, June 10–12, 2009. Proceedings, Part I, volume 5517 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02477-7 (print), 3-642-02478-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02478-8.

Chaudron:2008:CBS

[CSR08] Michel R. V. Chaudron, Clemens Szyperski, and Ralf Reussner, editors. Component-Based Software Engineering: 11th International Symposium, CBSE 2008, Karlsruhe, Germany, October 14–17, 2008. Proceedings, volume 5282 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87890-4 (print), 3-540-87891-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87891-9.

Coudert:2008:AHM

[CSRS08] David Coudert, David Simplot-Ryl, and Ivan Stojmenovic, editors. Ad-hoc, Mobile and Wireless Networks: 7th International Conference, ADHOC-NOW 2008 Sophia-Antipolis, France, September 10– 12, 2008 Proceedings, volume 5198 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85208-5 (print), 3-540-85209-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85209-4.

Crnkovic:2004:CBS

[CSSW04] Ivica Crnković, Judith A.

Stafford, Heinz W. Schmidt, and Kurt Wallnau, editors. Component-Based Software Engineering: 7th International Symposium, CBSE 2004, Edinburgh, UK, May 24– 25, 2004: Proceedings, volume 3054 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21998-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 C39 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3054.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21998-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3054;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97813.

Cazzola:2000:RSE

[CST00] Walter Cazzola, Robert J. Stroud, and Francisco Tisato. Reflection and Software Engineering, volume 1826 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67761-5. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 227 pp. LCCN QA76.758 .R44 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1826.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67761-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1826.

Coeurjolly:2008:DGC

[CSTD08] David Coeurjolly, Isabelle Sivignon, Laure Tougne, and Florent Dupont, editors. Discrete Geometry for Computer Imagery: 14th IAPR International Conference, DGCI 2008, Lyon, France, April 16– 18, 2008. Proceedings, volume 4992 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79125-6 (print), 3-540-79126-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79126-3.

Christiansen:2005:CSL

[CSV05] Henning Christiansen, Peter Rossen Skadhauge, and Jørgen Villadsen, editors. Constraint solving and language processing: first international workshop, CSLP 2004, Roskilde, Denmark, September 1– 3, 2004, Revised selected and invited papers, volume 3438 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26165-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        26165-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3438;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136253.

Cavalcanti:2006:RTS

[CSW06] Ana Cavalcanti, Augusto Sampaio, and Jim Woodcock, editors. Refinement Techniques in Software Engineering: First Pernambuco Summer School on Software Engineering, PSSE 2004, Recife, Brazil, November 23–December 5, 2004 Revised Lectures, volume 3167 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46253-8 (print), 3-540-46254-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46254-5.

Chung:2009:ISA

[CSY09] Kyo-Il Chung, Kiwook Sohn, and Moti Yung, editors. Information Security Applications: 9th International Workshop, WISA 2008, Jeju Island, Korea, September 23–25, 2008, Revised Selected Papers, volume 5379 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00305-2 (print), 3-642-00306-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00306-6.

Cyre:1999:CSS

[CT99] Walling Cyre and William M.

Tepfenhart, editors. Conceptual structures: standards and practices: 7th International Conference on Conceptual Structures, ICCS’99, Blacksburg, VA, USA, July 12–15, 1999: Proceedings, volume 1640 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66223-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1640; Q387.2 .I57 1999. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1640.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        66223-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1640.

Calzarossa:2002:PEC

[CT02] Maria Carla Calzarossa and Salvatore Tucci, editors. Performance Evaluation of Complex Systems: Techniques and Tools: Performance 2002. Tutorial Lectures, volume 2459 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44252-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.E9 P436 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2459.
                                        htm;   http://link.
                                        springer.de/link/service/
                                        series/0558/tocs/t2459.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        44252-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2459.

Cappelli:2003:AAA

[CT03] Amedeo Cappelli and Franco Turini, editors. AI*IA 2003: Advances in Artificial Intelligence: 8th Congress of the Italian Association for Artificial Intelligence, Pisa, Italy, September, 2003: Proceedings, volume 2829 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20119-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2829.htm;
 http://www.springerlink.
com/content/978-3-540-
20119-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2829;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13658.

Carswell:2006:WWG

[CT06] James D. Carswell and Taro Tezuka, editors. Web and Wireless Geographical Information Systems: 6th International Symposium, W2GIS 2006, Hong Kong, China, December 4–5, 2006. Proceedings, volume 4295 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49466-9 (print), 3-540-49467-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49467-6.

Chahed:2007:NCO

[CT07] Tijani Chahed and Bruno Tuffin, editors. Network Control and Optimization: First EuroFGI International Conference, NET-COOP 2007, Avignon, France, June 5– 7, 2007. Proceedings, volume 4465 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72708-6 (print), 3-540-72709-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72709-5.

Collet:2006:GPE

[CTE^{+}06] Pierre Collet, Marco Tomassini, Marc Ebner, Steven Gustafson, and Anikó Ekárt, editors. Genetic Programming: 9th European Conference, EuroGP 2006, Budapest, Hungary, April 10–12, 2006. Proceedings, volume 3905 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33143-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33143-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3905.

Corruble:2007:DSI

[CTS07] Vincent Corruble, Masayuki Takeda, and Einoshin Suzuki, editors. Discovery Science: 10th International Conference, DS 2007 Sendai, Japan, October 1–4, 2007. Proceedings, volume 4755 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75487-3 (print), 3-540-75488-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75488-6.

Chen:2010:WIS

[CTS10] Lei Chen, Peter Triantafillou, and Torsten Suel, editors. Web Information Systems Engineering — WISE 2010: 11th International Conference, Hong Kong, China, December 12– 14, 2010. Proceedings, volume 6488 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17615-1 (print), 3-642-17616-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17616-6.

Ceri:1993:DOO

[CTT93] Stefano Ceri, Katsumi Tanaka, and Shalom Tsur, editors. Deductive and object-oriented databases: third international conference, DOOD ’93, Phoenix, Arizona, USA, December 6–8, 1993: proceedings, volume 760 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57530-8 (Berlin), 0-387-57530-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55826 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0760.htm;
 http://www.springerlink.
com/content/978-3-540-

57530-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=760.

Chen:2010:WAI

[CTYG10] Lei Chen, Changjie Tang, Jun Yang, and Yunjun Gao, editors. Web-Age Information Management: 11th International Conference, WAIM 2010, Jiuzhaigou, China, July 15–17, 2010. Proceedings, volume 6184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14245-1 (print), 3-642-14246-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14246-8.

Claypool:2008:PAN

[CU08] Mark Claypool and Steve Uhlig, editors. Passive and Active Network Measurement: 9th International Conference, PAM 2008, Cleveland, OH, USA, April 29–30, 2008. Proceedings, volume 4979 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79231-7 (print), 3-540-79232-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79232-1.

Chaudhri:2002:XBD

[CUDL02] Akmal B. Chaudhri, Rainer Unland, Chabane Djeraba, and Wolfgang Lindner, editors. XML-Based Data Management and Multimedia Engineering—EDBT 2002: EDBT 2002 Workshops XMLDM, MDDE, and YRWS, Prague, Czech Republic, March 24–28, 2002. Revised Papers, volume 2490 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00130-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55844 2002b. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2490.htm;
 http://www.springerlink.
com/content/978-3-540-
00130-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2490.

This volume comprises papers from the following three workshops that were part of the complete program for the International Conference on Extending Database Technology (EDBT) … XML-Based Data Management (XMLDM), Second International Workshop on Multimedia Data and Document Engineering (MDDE), Young Researchers Workshop (YRWS).

Conejo:2004:CTA

[CUdlC04] Ricardo Conejo, Maite Urretavizcaya, and José-Luis Pérez de-la Cruz, editors. Current Topics in Artificial Intelligence: 10th Conference of the Spanish Association for Artificial Intelligence, CAEPIA 2003, and 5th Conference on Technology Transfer, TTIA 2003, San Sebastian, Spain, November 12–14, 2003: revised Selected Papers, volume 3040 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22218-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q335 .C87 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3040.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22218-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3040;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98369.

CuevaLovelle:2003:WEI

[Cue03] Juan Manuel Cueva Lovelle, editor. Web engineering: International Conference, ICWE 2003, Oviedo, Spain, July 14–18, 2003: Proceedings, volume 2722 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40522-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .I28 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2722.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40522-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2722.

Cuppens:2000:CSE

[Cup00] Frederic Cuppens, editor. Computer security — ESORICS 2000: 6th European Symposium on Research in Computer Security, Toulouse, France, October 4–6, 2000: Proceedings, volume 1895 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41031-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1895. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1895.htm;
 http://www.springerlink.
com/content/978-3-540-
41031-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1895.

Curien:2009:TLC

[Cur09] Pierre-Louis Curien, editor. Typed Lambda Calculi and Applications: 9th International Conference, TLCA 2009, Brasilia, Brazil, July 1–3, 2009. Proceedings, volume 5608 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02272-3 (print), 3-642-02273-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02273-9.

Chandru:1996:FST

[CV96] V. Chandru and V. Vinay, editors. Foundations of software technology and theoretical computer science: 16th conference, Hyderabad, India, December 18–20, 1996: proceedings, volume 1180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-62034-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F692 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1180.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62034-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1180.

“FST and TCS 16” “The proceedings of the Sixteenth Annual Conference on Foundations of Software Technology and Theoretical Computer Science”–Pref.

Camurri:2004:GBC

[CV04a] Antonio Camurri and Gualtiero Volpe, editors. Gesture-Based Communication in Human-Computer Interaction: 5th International Gesture Workshop, GW 2003: Genova, Italy, April 15–17, 2003: Selected Revised Papers, volume 2915 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21072-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 G473 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2915.
htm;   http://www.springerlink.
com/content/978-3-540-
21072-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2915;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95740.

Canteaut:2004:PCI

[CV04b] Anne Canteaut and Kapaleeswaran Viswanathan, editors. Progress in cryptology: INDOCRYPT 2004: 5th International Conference on Cryptology in India, Chennai, India, December 20–22, 2004: Proceedings, volume 3348 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24130-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I5535 2004.

                                        URL http://springerlink.
                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3348;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        24130-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3348;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104579.

Caires:2007:CCT

[CV07] Luís Caires and Vasco T. Vasconcelos, editors. CONCUR 2007 — Concurrency Theory: 18th International Conference, CONCUR 2007, Lisbon, Portugal, September 3–8, 2007. Proceedings, volume 4703 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74406-1 (print), 3-540-74407-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74407-8.

Caputo:2008:CVI

[CV08] Barbara Caputo and Markus Vincze, editors. Cognitive Vision: 4th International Workshop, ICVW 2008, Santorini, Greece, May 12, 2008, Revised Selected Papers, volume 5329 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92780-8 (print), 3-540-92781-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92781-5.

Clarke:2010:LPA

[CV10] Edmund M. Clarke and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 16th International Conference, LPAR-16, Dakar, Senegal, April 25–May 1, 2010, Revised Selected Papers, volume 6355 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17510-4 (print), 3-642-17511-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17511-4.

Cabot:2011:TPM

[CV11] Jordi Cabot and Eelco Visser, editors. Theory and Practice of Model Transformations: 4th International Conference, ICMT 2011, Zurich, Switzerland, June 27– 28, 2011. Proceedings, volume 6707 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21731-1 (print), 3-642-21732-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21732-6.

Csuhaj-Varju:2007:FCT

[CVÉ07] Erzsébet Csuhaj-Varjú and Zoltán Ésik, editors. Fundamentals of Computation Theory: 16th International Symposium, FCT 2007, Budapest, Hungary, August 27–30, 2007. Proceedings, volume 4639 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74239-5 (print), 3-540-74240-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74240-1.

Cotta:2007:ECC

[CvH07] Carlos Cotta and Jano van Hemert, editors. Evolutionary Computation in Combinatorial Optimization: 7th European Conference, EvoCOP 2007, Valencia, Spain, April 11– 13, 2007. Proceedings, volume 4446 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71614-9 (print), 3-540-71615-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71615-0.

Castano:2012:ACM

[CVLL12] Silvana Castano, Panos Vassiliadis, Laks V. Lakshmanan, and Mong Li Lee, editors. Advances in Conceptual Modeling: ER 2012 Workshops CMS, ECDM-NoCoDA, MoDIC, MORE-BI, RIGiM, SeCoGIS, WISM, Florence, Italy, October 15– 18, 2012. Proceedings, volume 7518 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33998-0 (print), 3-642-33999-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33999-8.

Cranefield:2012:COI

[CvRVSN12] Stephen Cranefield, M. Birna van Riemsdijk, Javier Vázquez-Salceda, and Pablo Noriega, editors. Coordination, Organizations, Institutions, and Norms in Agent System VII: COIN 2011 International Workshops, COIN@AAMAS 2011, Taipei, Taiwan, May 3, 2011, COIN@WI-IAT 2011, Lyon, France, August 22, 2011, Revised Selected Papers, volume 7254 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35544-7 (print), 3-642-35545-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35545-5.

Chi:2004:WCC

[CvSW04] Chi-Hung Chi, Maarten van Steen, and Craig Wills, editors. Web Content Caching and Distribution: 9th International Workshop, WCW 2004, Beijing, China, October 18–20, 2004. Proceedings, volume 3293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23516-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3293.htm;
http://www.springerlink.
com/content/978-3-540-
23516-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3293;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101692.

Caseli:2012:CPP

[CVTP12] Helena Caseli, Aline Villavicencio, António Teixeira, and Fernando Perdigão, editors. Computational Processing of the Portuguese Language: 10th International Conference, PRO-POR 2012, Coimbra, Portugal, April 17–20, 2012. Proceedings, volume 7243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28884-7 (print), 3-642-28885-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28885-2.

Cervesato:2008:LPA

[CVV08] Iliano Cervesato, Helmut Veith, and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 15th International Conference, LPAR 2008, Doha, Qatar, November 22–27, 2008. Proceedings, volume 5330 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89438-1 (print), 3-540-89439-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
89439-1.

Clarke:2012:PSI

[CVV12] Edmund Clarke, Irina Virbitskaite, and Andrei Voronkov, editors. Perspectives of Systems Informatics: 8th International Andrei Ershov Memorial Conference, PSI 2011, Novosibirsk, Russia, June 27–July 1, 2011, Revised Selected Papers, volume 7162 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29708-0 (print), 3-642-29709-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29709-0.

Cori:1988:SAS

[CW88] Robert Cori and Martin Wirsing, editors. STACS 88: 5th Annual Symposium on Theoretical Aspects of Computer Science, Bordeaux, France, February 11–13, 1988, proceedings, volume 294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 3-540-18834-7, 0-387-18834-7.

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.294. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0294.htm;
 http://www.springerlink.
com/content/978-3-540-
18834-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=294.

Cohen:1989:CTA

[CW89] G. Cohen and J. Wolfmann, editors. Coding theory and applications: 3rd international colloquium, Toulon, France, November 2–4, 1988: proceedings, volume 388 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51643-3 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .C67 1989. DM47.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0388.
htm;   http://www.springerlink.
com/content/978-0-387-
51643-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=388.

Proceedings of the colloquium “Trois journées sur le codage.”.

Castelfranchi:1994:ASS

[CW94a] Cristiano Castelfranchi and Eric Werner, editors. Artificial social systems: 4th European Workshop on Modelling Autonomous Agents in a Multi-Agent World, MAA-MAW ’92, S. Martino al Cimino, Italy, July 29–31, 1992: Selected Papers, volume 830 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58266-5 (Berlin), 0-387-58266-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HM131 .E87 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0830.htm;
http://www.springerlink.
com/content/978-3-540-
58266-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=830.

Chmora:1994:ECC

[CW94b] Andrew Chmora and Stephen B. Wicker, editors. Error control, cryptology, and speech compression: Workshop on Information Protection, Moscow, Russia, December 6–9, 1993: Selected Papers, volume 829 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58265-7 (Berlin), 0-387-58265-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 W668 1993. DM39.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0829.
htm;   http://www.springerlink.
com/content/978-3-540-
58265-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=829.

Ciancarini:1999:CLM

[CW99] Paolo Ciancarini and Alexander Wolf, editors. Coordination languages and models: Third International Conference, COORDINATION ’99, Amsterdam, The Netherlands, April 26–28, 1999: Proceedings, volume 1594 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65836-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1594.

URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t1594.
htm;   http://www.springerlink.
com/content/978-3-540-
65836-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1594.

Choueiry:2000:ARA

[CW00] Berthe Y. Choueiry and Toby Walsh, editors. Abstraction, reformulation, and approximation: 4th International Symposium, SARA 2000, Lake LBJ, Texas, USA, July 26–29, 2000: Proceedings, volume 1864 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67839-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1864.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1864.
htm;   http://www.springerlink.
com/content/978-3-540-
67839-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1864.

Ciancarini:2001:AOS

[CW01] Paolo Ciancarini and Michael J. Wooldridge, editors. Agent-oriented software engineering: First International Workshop, AOSE 2000, Limerick, Ireland, June 10, 2000: Revised Papers, volume 1957 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41594-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .A57 2000; QA267.A1 L43 no.1957. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1957.htm;
 http://www.springerlink.
com/content/978-3-540-
41594-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1957.

Clark:2002:OMO

[CW02] Tony Clark and Jos B. Warmer, editors. Object Modeling with the OCL: the rationale behind the Object Constraint Language, volume 2263 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43169-1. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 279 pp. LCCN QA76.9.O35 O242 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2263.htm;
http://www.springerlink.
com/content/978-3-540-
43169-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2263.

Conradi:2003:EMS

[CW03] Reidar Conradi and Alf Inge Wang, editors. Empirical Methods and Studies in Software Engineering: Experiences from ESER-NET, volume 2765 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40672-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .E52 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2765.htm;
http://www.springerlink.
com/content/978-3-540-
40672-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2765;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11962.

Ciancarini:2006:CML

[CW06] Paolo Ciancarini and Herbert Wiklicky, editors. Coordination Models and Languages: 8th International Conference, Coordination 2006, Bologna, Italy, June 14– 16, 2006. Proceedings, volume 4038 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34694-5 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34694-4.

Chen:2007:ACM

[CW07] Peter P. Chen and Leah Y. Wong, editors. Active Conceptual Modeling of Learning: Next Generation Learning-Base System Development, volume 4512 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77502-1 (print), 3-540-77503-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77503-4.

Chechik:2009:FAS

[CW09] Marsha Chechik and Martin Wirsing, editors. Fundamental Approaches to Software Engineering: 12th International Conference, FASE 2009, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2009, York, UK, March 22– 29, 2009. Proceedings, volume 5503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00592-6 (print), 3-642-00593-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00593-0.

Cezon:2011:TSB

[CW11] Michel Cezon and Yaron Wolfsthal, editors. Towards a Service-Based Internet. ServiceWave 2010 Workshops: International Workshops, OCS, EMSOA, SMART, and EDBPM 2010, Ghent, Belgium, December 13– 15, 2010, Revised Selected Papers, volume 6569 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22759-7 (print), 3-642-22760-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22760-8.

Chang:2007:AWN

[CWC^{+}07] Kevin Chen-Chuan Chang, Wei Wang, Lei Chen, Clarence A. Ellis, Ching-Hsien Hsu, Ah Chung Tsoi, and Haixun Wang, editors. Advances in Web and Network Technologies, and Information Management: APWeb/ WAIM 2007 International Workshops: DB-MAN 2007, WebETrends 2007, PAIS 2007 and ASWAN 2007, Huang Shan, China, June 16– 18, 2007. Proceedings, volume 4537 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72908-9 (print), 3-540-72909-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72909-9.

Chawla:2009:NFA

[CWiM^{+}09] Sanjay Chawla, Takashi Washio, Shin ichi Minato, Shusaku Tsumoto, Takashi Onoda, Seiji Yamada, and Akihiro Inokuchi, editors. New Frontiers in Applied Data Mining: PAKDD 2008 International Workshops, Osaka, Japan, May 20–23, 2008. Revised Selected Papers, volume 5433 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00398-2 (print), 3-642-00399-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00399-8.

Cheung:2001:AKD

[CWL01] David Cheung, Graham J. Williams, and Qing Li, editors. Advances in knowledge discovery and data mining: 5th Pacific-Asia Conference, PAKDD 2001, Hong Kong, China, April 16–18, 2001: Proceedings, volume 2035 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41910-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 P17 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2035.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41910-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2035.

Corchado:2009:HAI

[CWO^{+}09] Emilio Corchado, Xindong Wu, Erkki Oja, Álvaro Herrero, and Bruno Baruque, editors. Hybrid Artificial Intelligence Systems: 4th International Conference, HAIS 2009, Salamanca, Spain, June 10–12, 2009. Proceedings, volume 5572 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02318-5 (print), 3-642-02319-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02319-4.

Chen:2006:ISI

[CWY^{+}06] Hsinchun Chen, Fei Yue Wang, Christopher C. Yang, Daniel Zeng, Michael Chau, and Kuiyu Chang, editors. Intelligence and Security Informatics: International Workshop, WISI 2006, Singapore, April 9, 2006. Proceedings, volume 3917 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33361-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33361-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3917.

Chau:2012:ISI

[CWYC12] Michael Chau, G. Alan Wang, Wei Thoo Yue, and Hsinchun Chen, editors. Intelligence and Security Informatics: Pacific Asia Workshop, PAISI 2012, Kuala Lumpur, Malaysia, May 29, 2012. Proceedings, volume 7299 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30427-3 (print), 3-642-30428-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30428-6.

Chau:2011:ISI

[CWZ^{+}11a] Michael Chau, G. Alan Wang, Xiaolong Zheng, Hsinchun Chen, Daniel Zeng, and Wenji Mao, editors. Intelligence and Security Informatics: Pacific Asia Workshop, PAISI 2011, Beijing, China, July 9, 2011. Proceedings, volume 6749 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22038-X (print), 3-642-22039-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22039-5.

Chen:2011:BRA

[CWZ11b] Jianer Chen, Jianxin Wang, and Alexander Zelikovsky, editors. Bioinformatics Research and Applications: 7th International Symposium, IS-BRA 2011, Changsha, China, May 27–29, 2011. Proceedings, volume 6674 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21259-X (print), 3-642-21260-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21260-4.

Chae:2003:ISA

[CY03] Kijoon Chae and Moti Yung, editors. Information Security Applications: 4th International Workshop, WISA 2003, Jeju Island, Korea, August 25–27, 2003: Revised Papers, volume 2908 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20827-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 W577 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2908.
htm;   http://www.springerlink.
com/content/978-3-540-
20827-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2908;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95188.

Corchado:2009:IDE

[CY09] Emilio Corchado and Hujun Yin, editors. Intelligent Data Engineering and Automated Learning — IDEAL 2009: 10th International Conference, Burgos, Spain, September 23–26, 2009. Proceedings, volume 5788 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04393-3 (print), 3-642-04394-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04394-9.

Chen:2010:TSF

[CY10] Liqun Chen and Moti Yung, editors. Trusted Systems: First International Conference, IN-TRUST 2009, Beijing, China, December 17–19, 2009. Revised Selected Papers, volume 6163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14596-5 (print), 3-642-14597-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14597-1.

Chen:2011:TSS

[CY11a] Liqun Chen and Moti Yung, editors. Trusted Systems: Second International Conference, IN-TRUST 2010, Beijing, China, December 13–15, 2010, Revised Selected Papers, volume 6802 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25282-6 (print), 3-642-25283-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25283-9.

Chung:2011:ISA

[CY11b] Yongwha Chung and Moti Yung, editors. Information Security Applications: 11th International Workshop, WISA 2010, Jeju Island, Korea, August 24–26, 2010, Revised Selected Papers, volume 6513 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17954-1 (print), 3-642-17955-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17955-6.

Chim:2012:ICS

[CY12] Tat Wing Chim and Tsz Hon Yuen, editors. Information and Communications Security: 14th International Conference, ICICS 2012, Hong Kong, China, October 29–31, 2012. Proceedings, volume 7618 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34128-4 (print), 3-642-34129-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34129-8.

Corchado:2006:IDE

[CYBF06] Emilio Corchado, Hujun Yin, Vicente Botti, and Colin Fyfe, editors. Intelligent Data Engineering and Automated Learning — IDEAL 2006: 7th International Conference, Burgos, Spain, September 20–23, 2006. Proceedings, volume 4224 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45485-3 (print), 3-540-45487-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45487-8.

Chen:2009:ISI

[CYCL09] Hsinchun Chen, Christopher C. Yang, Michael Chau, and Shu-Hsing Li, editors. Intelligence and Security Informatics: Pacific Asia Workshop, PAISI 2009, Bangkok, Thailand, April 27, 2009. Proceedings, volume 5477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01392-9 (print), 3-642-01393-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01393-5.

Cao:2004:PDP

[CYGL04] Jiannong Cao, Laurence T. Yang, Minyi Guo, and Francis Lau, editors. Parallel and distributed processing and applications: second international symposium, ISPA 2004, Hong Kong, China, December 13– 15, 2004: Proceedings, volume 3358 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24128-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I88 2004. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3358;
 http://www.springerlink.
com/content/978-3-540-
24128-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3358.

Calero:2011:ATC

[CYM^{+}11] Jose M. Alcaraz Calero, Laurence T. Yang, Félix Gómez Mármol, Luis Javier García Villalba, Andy Xiaolin Li, and Yan Wang, editors. Autonomic and Trusted Computing: 8th International Conference, ATC 2011, Banff, Canada, September 2–4, 2011. Proceedings, volume 6906 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23495-X (print), 3-642-23496-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
23496-5.

Cortadella:2002:CHD

[CYR02] J. Cortadella, Alex Yakovlev, and Grzegorz Rozenberg, editors. Concurrency and hardware design: advances in Petri nets, volume 2549 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00199-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .C662 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2549.htm;
http://www.springerlink.
com/content/978-3-540-
00199-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2549.

Also available via the World Wide Web.

Chen:2012:TST

[CYZ12] Liqun Chen, Moti Yung, and Liehuang Zhu, editors. Trusted Systems: Third International Conference, INTRUST 2011, Beijing, China, November 27–29, 2011, Revised Selected Papers, volume 7222 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32297-2 (print), 3-642-32298-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32298-3.

Carle:2002:PHS

[CZ02] Georg Carle and Martina Zitterbart, editors. Protocols for High Speed Networks: 7th IFIP/IEEE International Workshop, PfHSN 2002, Berlin, Germany, April 22–24, 2002. Proceedings, volume 2334 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43658-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.55 .I3534 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2334.
htm;   http://www.springerlink.
com/content/978-3-540-
43658-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2334.

Chadwick:2005:PKI

[CZ05] David Chadwick and Gansen Zhao, editors. Public key infrastructure: Second European PKI Workshop: Research and Applications, EuroPKI 2005, Canterbury, UK, June 30–July 1, 2005: revised selected papers, volume 3545 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28062-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E976 2005.

URL http://www.springerlink.
com/content/978-3-540-
28062-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3545.

Cao:2010:ADMc

[CZF10] Longbing Cao, Jiang Zhong, and Yong Feng, editors. Advanced Data Mining and Applications: 6th International Conference, ADMA 2010, Chongqing, China, November 19–21, 2010, Proceedings, Part II, volume 6441 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17312-8 (print), 3-642-17313-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17313-4.

Chapman:2008:PPM

[CZG^{+}08] Barbara Chapman, Weiming Zheng, Guang R. Gao, Mitsuhisa Sato, Eduard Ayguadé, and Dongsheng Wang, editors. A Practical Programming Model for the Multi-Core Era: 3rd International Workshop on OpenMP, IWOMP 2007, Beijing, China, June 3–7, 2007 Proceedings, volume 4935 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69302-5 (print), 3-540-69303-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69303-1.

Chen:2008:STP

[CZZ08] Qingfeng Chen, Chengqi Zhang, and Shichao Zhang, editors. Secure Transaction Protocol Analysis: Models and Applications, volume 5111 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85073-2 (print), 3-540-85074-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 221 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85074-8.

Dumke:2001:NAS

[DA01] Reiner Dumke and Alain Abran, editors. New approaches in software measurement: 10th International Workshop, IWSM 2000, Berlin, Germany, October 4– 6, 2000: Proceedings, volume 2006 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41727-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2006; QA76.76.S65 I97 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2006.htm;
http://www.springerlink.
com/content/978-3-540-
41727-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2006.

Domingue:2008:SWA

[DA08] John Domingue and Chutiporn Anutariya, editors. The Semantic Web: 3rd Asian Semantic Web Conference, ASWC 2008, Bangkok, Thailand, December 8– 11, 2008. Proceedings., volume 5367 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89703-8 (print), 3-540-89704-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89704-0.

deAlfaro:2009:FSS

[dA09] Luca de Alfaro, editor. Foundations of Software Science and Computational Structures: 12th International Conference, FOSSACS 2009, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2009, York, UK, March 22– 29, 2009. Proceedings, volume 5504 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00595-0 (print), 3-642-00596-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00596-1.

DeAntonio:2001:IVA

[DAB01] Angelica De Antonio, Ruth Aylett, and Daniel Ballin, editors. Intelligent virtual agents: Third International Workshop, IVA 2001, Madrid, Spain, September 10– 11, 2001: Proceedings, volume 2190 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42570-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65 I93 2001; QA267.A1 L43 no.2190. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2190.htm;
http://www.springerlink.
com/content/978-3-540-
42570-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2190.

DiChio:2012:AEC

[DAC^{+}12] Cecilia Di Chio, Alexandros Agapitos, Stefano Cagnoni, Carlos Cotta, Francisco Fernández de Vega, Gianni A. Di Caro, Rolf Drechsler, Anikó Ekárt, Anna I. Esparcia-Alcázar, Muddassar Farooq, William B. Langdon, Juan J. Merelo-Guervós, Mike Preuss, Hendrik Richter, Sara Silva, Anabela Simões, Giovanni Squillero, Ernesto Tarantino, Andrea G. B. Tettamanzi, Julian Togelius, Neil Urquhart, A. Şima Uyar, and Georgios N. Yannakakis, editors. Applications of Evolutionary Computation: EvoApplications 2012: EvoCOM-NET, EvoCOMPLEX, EvoFIN, EvoGAMES, EvoHOT, EvoIASP, EvoNUM, EvoPAR, EvoRISK, EvoS-TIM, and EvoSTOC, Málaga, Spain, April 11–13, 2012, Proceedings, volume 7248 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29177-5 (print), 3-642-29178-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29178-4.

Daduna:2001:QND

[Dad01] Hans Daduna. Queueing Networks with Discrete Time Scale: Explicit Expressions for the Steady State Behavior of Discrete Time Stochastic Networks, volume 2046 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42357-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 138 pp. LCCN QA274.8 .D33 2001; QA267.A1 L43 no.2046.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2046.
htm;   http://www.springerlink.
com/content/978-3-540-
42357-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2046.

Dainoff:2007:EHA

[Dai07] Marvin J. Dainoff, editor. Ergonomics and Health Aspects of Work with Computers: International Conference, EHAWC 2007, Held as Part of HCI International 2007, Beijing, China, July 22–27, 2007. Proceedings, volume 4566 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73332-9 (print), 3-540-73333-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73333-1.

Dale:1992:AAN

[Dal92] Robert Dale, editor. Aspects of automated natural language generation: 6th International Workshop on Natural Language Generation, Trento, Italy, April 5–7, 1992: proceedings, volume 587 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 0-387-55399-1 (USA), 3-540-55399-1 (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 A86 1992.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0587.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        55399-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=587.

Damgaard:1991:ACE

[Dam91] I. B. Damgård, editor. Advances in cryptology — EUROCRYPT ’90: Workshop on the Theory and Application of Cryptographic Techniques, Aarhus, Denmark, May 21–24, 1990: proceedings, volume 473 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-53587-X (New York), 3-540-53587-X (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 1990. DM69.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0473.
htm;   http://www.springerlink.
com/content/978-0-387-
53587-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=473.

Dam:1997:AVM

[Dam97] Mads Dam, editor. Analysis and verification of multiple-agent languages: 5th LOMAPS Workshop, Stockholm, Sweden, June 24–26, 1996: Selected Papers, volume 1192 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62503-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .L662 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1192.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62503-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1192.

Damgaard:1999:LDS

[Dam99a] I. B. (Ivan Bjerre) Damgård, editor. Lectures on data security: modern cryptology in theory and practice, volume 1561 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65757-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25L46 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1561.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65757-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1561.

Dams:1999:TPA

[Dam99b] Dennis Dams, editor. Theoretical and practical aspects of SPIN model checking: 5th and 6th international SPIN workshops, Trento, Italy, July 5, 1999, Toulouse, France, September 21 and 24, 1999: Proceedings, volume 1680 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66499-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 S68 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1680.
htm;   http://www.springerlink.
com/content/978-3-540-
66499-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1680.

Danezis:2012:FCDb

[Dan12] George Danezis, editor. Financial Cryptography and Data Security: 15th International Conference, FC 2011, Gros Islet, St. Lucia, February 28 — March 4, 2011, Revised Selected Papers, volume 7035 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27575-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27575-3.

Darnell:1997:CCI

[Dar97] Mike Darnell, editor. Cryptography and coding: 6th IMA conference, Cirencester, UK, December 17–19, 1997: proceedings, volume 1355 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63927-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .C76 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1355.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63927-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1355.

Datta:2009:ACS

[Dat09] Anupam Datta, editor. Advances in Computer Science — ASIAN 2009. Information Security and Privacy: 13th Asian Computing Science Conference, Seoul, Korea, December 14–16, 2009. Proceedings, volume 5913 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10621-8 (print), 3-642-10622-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10622-4.

Dau:2003:LSC

[Dau03] Frithjof Dau, editor. The Logic System of Concept Graphs with Negation: and its relationship to predicate logic, volume 2892 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20607-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2892.htm;
http://www.springerlink.
com/content/978-3-540-
20607-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2892;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94030.

Davenport:1981:IAF

[Dav81] James Harold Davenport. On the integration of algebraic functions, volume 102 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10290-6 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). 197 pp. LCCN QA341 .D33. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0102.htm;
 http://www.springerlink.
com/content/978-0-387-
10290-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=102.

Davenport:1989:EEC

[Dav89] James Harold Davenport, editor. EUROCAL ’87: European Conference on Computer Algebra, Leipzig, GDR, June 2–5, 1987: proceedings, volume 378 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51517-8 (New York), 3-540-51517-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E861 1987. US$39.40. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0378.
htm;   http://www.springerlink.
com/content/978-0-387-
51517-5;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=378.

Davies:1991:ACE

[Dav91] Donald Watts Davies, editor. Advances in cryptology — EUROCRYPT ’91: Workshop on the Theory and Application of Cryptography Techniques, Brighton, UK, April 8–11, 1991: proceedings, volume 547 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54620-0 (New York), 3-540-54620-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0547.htm;
http://www.springerlink.
com/content/978-0-387-
54620-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=547.

Rev. and expanded papers from the meeting which was sponsored by the International Association for Cryptology Research (IACR) and others.

deBakker:1992:RTT

[dB92] J. W. de Bakker, editor. Real-time, theory in practice: REX Workshop, Mook, the Netherlands, June 3–7, 1991: proceedings, volume 600 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55564-1 (Berlin), 0-387-55564-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54 .R49 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0600.htm;
 http://www.springerlink.
com/content/978-3-540-
55564-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=600.

deBerg:1993:RSD

[dB93] Mark de Berg. Ray shooting, depth orders and hidden surface removal, volume 703 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57020-9 (Berlin), 0-387-57020-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 201 pp. LCCN QA447 .B47 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0703.htm;

http://www.springerlink.
com/content/978-3-540-
57020-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=703.

Revision of the author’s thesis (Ph. D.).

Dawson:1998:ISP

[DB98a] Ed (Edward) Dawson and Colin Boyd, editors. Information security and privacy: Third Australasian Conference, ACISP’98, Brisbane, Australia, July 13–15, 1998: Proceedings, volume 1438 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64732-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 A2792 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1438.
htm;   http://www.springerlink.
com/content/978-3-540-
64732-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1438.

Demeyer:1998:OOT

[DB98b] Serge Demeyer and Jan Bosch, editors. Object Oriented Technology— ECOOP’98 Workshop Reader: ECOOP’98 Workshops, Demos, and Posters, Brussels, Belgium, July 20–24, 1998. Proceedings, volume 1543 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65460-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35O27 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1543.htm;
 http://www.springerlink.
com/content/978-3-540-
65460-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1543.

Das:2002:DCM

[DB02] Sajal K. Das and Swapan Bhattacharya, editors. Distributed computing: mobile and wireless computing: 4th International Workshop, IWDC 2002, Calcutta, India, December 28–31, 2002: Proceedings, volume 2571 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00355-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I68 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2571.htm;
 http://www.springerlink.
com/content/978-3-540-
00355-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2571.

Also available via the World Wide Web.

Dean:2004:TFM

[DB04] C. Neville Dean and Raymond T. Boute, editors. Teaching Formal Methods: COLOG NET/FME Symposium, TFM 2004, Ghent, Belgium, November 18–19, 2004: Proceedings, volume 3294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23611-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3294.htm;
 http://www.springerlink.
com/content/978-3-540-
23611-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3294;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b102075.

Doherty:2007:ISD

[DB07] Gavin Doherty and Ann Blandford, editors. Interactive Systems. Design, Specification, and Verification: 13th International Workshop, DSVIS 2006, Dublin, Ireland, July 26–28, 2006. Revised Papers, volume 4323 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69553-2 (print), 3-540-69554-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69554-7.

DiNoia:2009:CWT

[DB09] Tommaso Di Noia and Francesco Buccafurri, editors. E-Commerce and Web Technologies: 10th International Conference, EC-Web 2009, Linz, Austria, September 1–4, 2009. Proceedings, volume 5692 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03963-4 (print), 3-642-03964-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03964-5.

Dorigo:2004:ACO

[DBB^{+}04] Marco Dorigo, Mauro Birattari, Christian Blum, Luca M. Gambardella, Francesco Mondada, and Thomas Stützle, editors. Ant colony, optimization and swarm intelligence: 4th international workshop, Ants 2004, Brussels, Belgium, September 5–8, 2004, proceedings, volume 3172 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22672-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
22672-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3172;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99492.

Dengel:2008:KAA

[DBB^{+}08a] Andreas R. Dengel, Karsten Berns, Thomas M. Breuel, Frank Bomarius, and Thomas R.

Roth-Berghofer, editors. KI 2008: Advances in Artificial Intelligence: 31st Annual German Conference on AI, KI 2008, Kaiserslautern, Germany, September 23– 26, 2008. Proceedings, volume 5243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85844-X (print), 3-540-85845-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85845-4.

Dorigo:2008:ACO

[DBB^{+}08b] Marco Dorigo, Mauro Birattari, Christian Blum, Maurice Clerc, Thomas Stützle, and Alan F. T. Winfield, editors. Ant Colony Optimization and Swarm Intelligence: 6th International Conference, ANTS 2008, Brussels, Belgium, September 22–24, 2008. Proceedings, volume 5217 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87526-3 (print), 3-540-87527-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87527-7.

Dumke:2008:SPP

[DBB^{+}08c] Reiner R. Dumke, René Braungarten, Günter Büren, Alain Abran, and Juan J. Cuadrado-Gallego, editors. Software Process and Product Measurement: International Conferences IWSM 2008, Metrikon 2008, and Mensura 2008 Munich, Germany, November 18–19, 2008. Proceedings, volume 5338 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89402-0 (print), 3-540-89403-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89403-2.

Dennis:2012:PMA

[DBB12a] Louise Dennis, Olivier Boissier, and Rafael H. Bordini, editors. Programming Multi-Agent Systems: 9th International Workshop, Pro-MAS 2011, Taipei, Taiwan, May 3, 2011, Revised Selected Papers, volume 7217 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31914-9 (print), 3-642-31915-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31915-0.

Dorigo:2012:SII

[DBB^{+}12b] Marco Dorigo, Mauro Birattari, Christian Blum, Anders Lyhne Christensen, Andries P. Engelbrecht, Roderich Groß, and Thomas Stützle, editors. Swarm Intelligence: 8th International Conference, ANTS 2012, Brussels, Belgium, September 12–14, 2012. Proceedings, volume 7461 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32649-8 (print), 3-642-32650-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32650-9.

deBra:2002:AHA

[dBBC02] Paul M. E. de Bra, Peter Brusilovsky, and Ricardo Conejo, editors. Adaptive hypermedia and adaptive Web-based systems: Second International Conference, AH 2002, Malaga, Spain, May 29–31, 2002: Proceedings, volume 2347 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43737-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I59 A36 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2347.htm;
 http://www.springerlink.
com/content/978-3-540-
43737-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2347.

deBoer:2004:FMC

[dBBG04] Frank S. de Boer, Marcello M. Bonsangue, and Susanne Graf, editors. Formal Methods for Components and Objects: Second International Symposium, FMCO 2003, Leiden, The Netherlands, November 4–7, 2003. revised lectures, volume 3188 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22942-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3188.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22942-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3188;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100112.

deBoer:2003:FMC

[dBBGdR03] Frank S. de Boer, Marcello M. Bonsangue, Susanne Graf, and Willem-Paul de Roever, editors. Formal Methods for Components and Objects: First International Symposium, FMCO 2002, Leiden, The Netherlands, November 5–8, 2002: Revised Lectures, volume 2852 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20303-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 F53 2002.

                                        URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t2852.
htm;   http://www.springerlink.
com/content/978-3-540-
20303-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2852;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b14033.

deBoer:2005:FMC

[dBBGdR05] Frank S. de Boer, Marcello M. Bonsangue, Susanne Graf, and Willem-Paul de Roever, editors. Formal methods for components and objects: third international symposium, FMCO 2004, Leiden, The Netherlands, November 2–5, 2004, revised lectures, volume 3657 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29131-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29131-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3657.

deBoer:2006:FMC

[dBBGdR06] Frank S. de Boer, Marcello M. Bonsangue, Susanne Graf, and Willem-Paul de Roever, editors. Formal Methods for Components and Objects: 4th International Symposium, FMCO 2005, Amsterdam, The Netherlands, November 1– 4, 2005, Revised Lectures, volume 4111 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36749-7 (print), 3-540-36750-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36750-5.

deBoer:2007:FMC

[dBBGdR07] Frank S. de Boer, Marcello M. Bonsangue, Susanne Graf, and Willem-Paul de Roever, editors. Formal Methods for Components and Objects: 5th International Symposium, FMCO 2006, Amsterdam, The Netherlands, November 7– 10, 2006, Revised Lectures, volume 4709 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74791-5 (print), 3-540-74792-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74792-5.

deBoer:2008:FMC

[dBBGdR08] Frank S. de Boer, Marcello M. Bonsangue, Susanne Graf, and Willem-Paul de Roever, editors. Formal Methods for Components and Objects: 6th International Symposium, FMCO 2007, Amsterdam, The Netherlands, October 24–26, 2007, Revised Lectures, volume 5382 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92187-7 (print), 3-540-92188-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92188-2.

deBoer:2010:FMC

[dBBHL10] Frank S. de Boer, Marcello M. Bonsangue, Stefan Hallerstede, and Michael Leuschel, editors. Formal Methods for Components and Objects: 8th International Symposium, FMCO 2009, Eindhoven, The Netherlands, November 4–6, 2009. Revised Selected Papers, volume 6286 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17070-6 (print), 3-642-17071-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17071-3.

deBoer:2009:FMC

[dBBM09] Frank S. de Boer, Marcello M. Bonsangue, and Eric Madelaine, editors. Formal Methods for Components and Objects: 7th International Symposium, FMCO 2008, Sophia Antipolis, France, October 21–23, 2008, Revised Lectures, volume 5751 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04166-3 (print), 3-642-04167-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04167-9.

Deb:2010:SEL

[DBC^{+}10] Kalyanmoy Deb, Arnab Bhattacharya, Nirupam Chakraborti, Partha Chakroborty, Swagatam Das, Joydeep Dutta, Santosh K. Gupta, Ashu Jain, Varun Aggarwal, Jürgen Branke, Sushil J. Louis, and Kay Chen Tan, editors. Simulated Evolution and Learning: 8th International Conference, SEAL 2010, Kanpur, India, December 1–4, 2010. Proceedings, volume 6457 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17297-0 (print), 3-642-17298-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17298-4.

DiChio:2010:AECb

[DBD^{+}10a] Cecilia Di Chio, Anthony Brabazon, Gianni A. Di Caro, Marc Ebner, Muddassar Farooq, Andreas Fink, Jörn Grahl, Gary Greenfield, Penousal Machado, Michael O’Neill, Ernesto Tarantino, and Neil Urquhart, editors. Applications of Evolutionary Computation: EvoApplications 2010: Evo-COMNET, EvoENVI-RONMENT, EvoFIN, EvoMUSART, and Evo-TRANSLOG, Istanbul, Turkey, April 7–9, 2010, Proceedings, Part II, volume 6025 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12241-8 (print), 3-642-12242-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12242-2.

Dorigo:2010:SII

[DBD^{+}10b] Marco Dorigo, Mauro Birattari, Gianni A. Di Caro, René Doursat, Andries P. Engelbrecht, Dario Floreano, Luca Maria Gambardella, Roderich Groß, Erol Şahin, Hiroki Sayama, and Thomas Stützle, editors. Swarm Intelligence: 7th International Conference, ANTS 2010, Brussels, Belgium, September 8–10, 2010. Proceedings, volume 6234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15460-3 (print), 3-642-15461-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15461-4.

DiChio:2011:AECb

[DBD^{+}11] Cecilia Di Chio, Anthony Brabazon, Gianni A. Di Caro, Rolf Drechsler, Muddassar Farooq, Jörn Grahl, Gary Greenfield, Christian Prins, Juan Romero, Giovanni Squillero, Ernesto Tarantino, Andrea G. B. Tettamanzi, Neil Urquhart, and A. Şima Uyar, editors. Applications of Evolutionary Computation: EvoApplications 2011: EvoCOMNET, EvoFIN, EvoHOT, Evo-MUSART, EvoSTIM, and EvoTRANSLOG, Torino, Italy, April 27– 29, 2011, Proceedings, Part II, volume 6625 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20519-4 (print), 3-642-20520-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20520-0.

deBakker:1989:LTB

[dBdRR89] J. W. de Bakker, W.- P. de Roever, and Grzegorz Rozenberg, editors. Linear time, branching time, and partial order in logics and models for concurrency: School/Workshop, Noordwijkerhout, The Netherlands, May 30–June 3, 1988, volume 354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51080-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .L5231 1989. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0354.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51080-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=354.

“Based on the School/Workshop on Linear Time, Branching Time, and Partial Order in Logics and Models for Concurrency, organized by the editors”– Pref.

deBakker:1990:SRD

[dBdRR90] J. W. de Bakker, W.-P. de Roever, and Grzegorz Rozenberg, editors. Stepwise refinement of distributed systems: models, formalisms, correctness: REX workshop, Mook, the Netherlands, May 29–June 2, 1989: proceedings, volume 430 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52559-9 (Berlin), 0-387-52559-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 S75 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0430.htm;
 http://www.springerlink.
com/content/978-3-540-
52559-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=430.

deBakker:1991:FOO

[dBdRR91] J. W. de Bakker, W.- P. de Roever, and Grzegorz Rozenberg. Foundations of object-oriented languages: REX School/ Workshop, Noordwijkerhout, the Netherlands, May 28–June 1, 1990: proceedings, volume 489 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-53931-X (Berlin), 0-387-53931-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 442 pp. LCCN QA76.64 .R49 1991. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0489.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        53931-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=489.

deBakker:1993:SFA

[dBdRR93] J. W. de Bakker, W.- P. de Roever, and Grzegorz Rozenberg, editors. Semantics: foundations and applications: REX Workshop, Beekbergen, the Netherlands, June 1–4, 1992: proceedings, volume 666 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56596-5 (Berlin), 0-387-56596-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .R49 1992. DM122.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0666.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        56596-3;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=666.

deBakker:1994:DCR

[dBdRR94] J. W. de Bakker, W.- P. de Roever, and Grzegorz Rozenberg, editors. A Decade of concurrency: reflections and perspectives: REX school/symposium, Noordwijkerhout, the Netherlands, June 1–4, 1993: proceedings, volume 803 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58043-3 (Berlin), 0-387-58043-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .D43 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0803.htm;
http://www.springerlink.
com/content/978-3-540-
58043-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=803.

Dorffner:2001:ANN

[DBH01] Georg Dorffner, Horst Bischof, and Kurt Hornik, editors. Artificial neural networks — ICANN 2001: International Conference, Vienna, Austria, August 21–25, 2001: Proceedings, volume 2130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42486-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2130. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2130.htm;
 http://www.springerlink.
com/content/978-3-540-
42486-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2130.

Dillmann:2010:KAA

[DBHS10] Rüdiger Dillmann, Jürgen Beyerer, Uwe D. Hanebeck, and Tanja Schultz, editors. KI 2010: Advances in Artificial Intelligence: 33rd Annual German Conference on AI, Karlsruhe, Germany, September 21–24, 2010. Proceedings, volume 6359 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16110-3 (print), 3-642-16111-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16111-7.

Devi:2009:APA

[DBM09] Sobha Lalitha Devi, António Branco, and Ruslan Mitkov, editors. Anaphora Processing and Applications: 7th Discourse Anaphora and Anaphor Resolution Colloquium, DAARC 2009 Goa, India, November 5–6, 2009 Proceedings, volume 5847 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04974-5 (print), 3-642-04975-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04975-0.

deBerg:2010:AEAa

[dBM10a] Mark de Berg and Ulrich Meyer, editors. Algorithms — ESA 2010: 18th Annual European Symposium, Liverpool, UK, September 6–8, 2010. Proceedings, Part I, volume 6346 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15774-2 (print), 3-642-15775-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15775-2.

deBerg:2010:AEAb

[dBM10b] Mark de Berg and Ulrich Meyer, editors. Algorithms — ESA 2010: 18th Annual European Symposium, Liverpool, UK, September 6–8, 2010. Proceedings, Part II, volume 6347 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15780-7 (print), 3-642-15781-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15781-3.

deBakker:1987:PPAa

[dBNT87a] J. W. (Jacobus Wilhelm) de Bakker, L. Nijman, and P. C. Treleaven, editors. PARLE, Parallel Architectures and Languages Europe: Eindhoven, The Netherlands, June 15–19, 1987: Proceedings. Part I. Parallel Architectures, volume 258 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-17943-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .C61926 1987. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0258.
htm;   http://www.springerlink.
com/content/978-0-387-
17943-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=258.

deBakker:1987:PPAb

[dBNT87b] J. W. (Jacobus Wilhelm)

de Bakker, L. Nijman, and P. C. Treleaven, editors. PARLE, Parallel Architectures and Languages Europe: Eindhoven, The Netherlands, June 15–19, 1987: Proceedings. Part II. Parallel Languages, volume 259 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-17945-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .C61926 1987. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0259.
htm;   http://www.springerlink.
com/content/978-0-387-
17945-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=259.

Deng:2005:ISP

[DBPZ05] Robert H. Deng, Feng Bao, HweeHwa Pang, and Jianying Zhou, editors. Information Security Practice and Experience: First International Conference, IS-PEC 2005, Singapore, April 11–14, 2005. Proceedings, volume 3439 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25584-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25584-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3439;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b107167.

deBakker:1986:CTC

[dBRdR86] J. W. de Bakker, Grzegorz Rozenberg, and W.- P. de Roever, editors. Current trends in concurrency: overviews and tutorials, volume 224 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16488-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .C7951 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0224.htm;
 http://www.springerlink.
com/content/978-0-387-
16488-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=224.

Proceedings of the ES-PRIT/LPC Advanced School on Current Trends in Concurrency, held June 10–21, 1985, at Noordwijkerhout, The Netherlands; sponsored by ESPRIT Project 415, Parallel Architectures and Languages for AIP/a VLSI directed approach.

Dignum:2009:AGS

[DBSvD09] Frank Dignum, Jeff Bradshaw, Barry Silverman, and Willem van Doesburg, editors. Agents for Games and Simulations: Trends in Techniques, Concepts and Design, volume 5920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-11197-1 (print), 3-642-11198-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11198-3.

Dieng:2000:KEK

[DC00] Rose Dieng and Olivier Corby, editors. Knowledge engineering and knowledge management: methods, models, and tools: 12th International Conference, EKAW 2000, Juan-les-Pins, France, October 2–6, 2000: Proceedings, volume 1937 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41119-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E92 2000; QA267.A1 L43 no.1937. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1937.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41119-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1937.

Dignum:2001:AMEb

[DC01] Frank Dignum and Ulisis Cortés. Agentmediated electronic commerce III: current issues in agent-based electronic commerce systems, volume 2003 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41749-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 191 pp. LCCN HF5548.32 .A34 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2003.htm;
 http://www.springerlink.
com/content/978-3-540-
41749-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2003.

DeDecker:2012:CMS

[DC12] Bart De Decker and David W. Chadwick, editors. Communications and Multimedia Security: 13th IFIP TC 6/TC 11 International Conference, CMS 2012, Canterbury, UK, September 3–5, 2012. Proceedings, volume 7394 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32804-0 (print), 3-642-32805-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32805-3.

DiChio:2010:AECa

[DCC^{+}10] Cecilia Di Chio, Stefano Cagnoni, Carlos Cotta, Marc Ebner, Anikó Ekárt, Anna I. Esparcia-Alcazar, Chi-Keong Goh, Juan J. Merelo, Ferrante Neri, Mike Preuß, Julian Togelius, and Georgios N. Yannakakis, editors. Applications of Evolutionary Computation: EvoApplications 2010: EvoCOMPLEX, EvoGAMES, EvoIASP, EvoINTELLIGENCE, EvoNUM, and EvoSTOC, Istanbul, Turkey, April 7–9, 2010, Proceedings, Part I, volume 6024 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12238-8 (print), 3-642-12239-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12239-2.

DiChio:2011:AECa

[DCC^{+}11] Cecilia Di Chio, Stefano Cagnoni, Carlos Cotta, Marc Ebner, Anikó Ekárt, Anna I. Esparcia-Alcázar, Juan J. Merelo, Ferrante Neri, Mike Preuss, Hendrik Richter, Julian Togelius, and Georgios N. Yannakakis, editors. Applications of Evolutionary Computation: EvoApplications 2011: EvoCOMPLEX, EvoGAMES, EvoIASP, EvoINTELLIGENCE, EvoNUM, and EvoSTOC, Torino, Italy, April 27–29, 2011, Proceedings, Part I, volume 6624 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20524-0 (print), 3-642-20525-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20525-5.

daCosta:2008:URS

[dCdF^{+}08] Paulo Cesar G. da Costa, Claudia d’Amato, Nicola Fanizzi, Kathryn B. Laskey, Kenneth J. Laskey, Thomas Lukasiewicz, Matthias Nickles, and Michael Pool, editors. Uncertainty Reasoning for the Semantic Web I: ISWC International Workshops, URSW 2005-2007, Revised Selected and Invited Papers, volume 5327 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89764-X (print), 3-540-89765-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89765-1.

Dawant:2012:BIR

[DCFR12] Benoît M. Dawant, Gary E. Christensen, J. Michael Fitzpatrick, and Daniel Rueckert, editors. Biomedical Image Registration: 5th International Workshop, WBIR 2012, Nashville, TN, USA, July 7–8, 2012. Proceedings, volume 7359 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31339-6 (print), 3-642-31340-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31340-0.

Dolev:2010:SSS

[DCFY10] Shlomi Dolev, Jorge Cobb, Michael Fischer, and Moti Yung, editors. Stabilization, Safety, and Security of Distributed Systems: 12th International Symposium, SSS 2010, New York, NY, USA, September 20– 22, 2010. Proceedings, volume 6366 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16022-0 (print), 3-642-16023-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16023-3.

delCerro:2012:LAI

[dCHM12] Luis Fariñas del Cerro, Andreas Herzig, and Jérôme Mengin, editors. Logics in Artificial Intelligence: 13th European Conference, JELIA 2012, Toulouse, France, September 26–28, 2012. Proceedings, volume 7519 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33352-4 (print), 3-642-33353-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33353-8.

Junior:2012:PLB

[dCJB12] Francisco Heron de Carvalho Junior and Luis Soares Barbosa, editors. Programming Languages: 16th Brazilian Symposium, SBLP 2012, Natal, Brazil, September 23–28, 2012. Proceedings, volume 7554 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33181-5 (print), 3-642-33182-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33182-4.

Dezani-Ciancaglini:1982:ISP

[DCM82] M. Dezani-Ciancaglini and Ugo Montanari, editors. International Symposium on Programming: 5th colloquium, Turin, April 6–8, 1982: proceedings, volume 137 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 3-540-11494-7 (Berlin), 0-387-11494-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I575.1982.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0137.
htm;   http://www.springerlink.
com/content/978-3-540-
11494-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=137.

Dillon:2009:AWS

[DCMS09] Tharam S. Dillon, Elizabeth Chang, Robert Meersman, and Katia Sycara, editors. Advances in Web Semantics I: Ontologies, Web Services and Applied Semantic Web, volume 4891 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-89783-6 (print), 3-540-89784-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89784-2.

Dezani-Ciancaglini:1995:TLC

[DCP95] M. Dezani-Ciancaglini and G. Plotkin, editors. Typed lambda calculi and applications: second International Conference on Typed Lambda Calculi and Applications, TLCA ’95, Edinburgh, United Kingdom, April 10–12, 1995: proceedings, volume 902 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59048-X (Berlin: =bk.), 0-387-59048-X (New York: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.5 .I58 1995. DM86.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0902.htm;
 http://www.springerlink.
com/content/978-3-540-
59048-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=902.

deCastro:2007:AIS

[dCVK07] Leandro Nunes de Castro, Fernando José Von Zuben, and Helder Knidel, editors. Artificial Immune Systems: 6th International Conference, ICARIS 2007, Santos, Brazil, August 26–29, 2007. Proceedings, volume 4628 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73921-1 (print), 3-540-73922-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73922-7.

DeMichelis:1995:ATP

[DD95] Giorgio De Michelis and Michel Diaz, editors. Application and theory of Petri nets, 1995: 16th International Conference, Turin, Italy, June 26–30, 1995: proceedings, volume 935 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60029-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A66 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0935.htm;
http://www.springerlink.
com/content/978-3-540-
60029-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=935.

Application and theory of Petri nets; 16th, 1995.

Danthine:1997:MSN

[DD97] A. Danthine and Christophe Diot, editors. From multimedia services to network services: 4th International COST 237 Workshop, Lisboa, Portugal, December 15– 19, 1997: proceedings, volume 1356 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63935-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I65 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1356.htm;
 http://www.springerlink.
com/content/978-3-540-
63935-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1356.

Das:2003:DCI

[DD03] Samir R. Das and Sajal K. Das, editors. Distributed Computing— IWDC 2003: 5th International Workshop, Kolkata, India, December 27–30, 2003: Proceedings, volume 2918 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20745-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I68 2003.

URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t2918.
htm;   http://www.springerlink.
com/content/978-3-540-
20745-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2918;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94926.

Diekert:2005:SAS

[DD05] Volker Diekert and Bruno Durand, editors. STACS 2005: 22nd annual symposium on theoretical aspects of computer science, Stuttgart, Germany, February 24– 26, 2005, proceedings, volume 3404 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24998-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.7 .S958 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24998-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3404;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106485.

Deng:2006:ACI

[DD06] Xiaotie Deng and Dingzhu Du, editors. Algorithms and computation: 16th international symposium, ISAAC 2005, Sanya, Hainan, China, December 19–21, 2005, proceedings, volume 3827 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30935-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        30935-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3827.

Dietrich:2007:FCD

[DD07] Sven Dietrich and Rachna Dhamija, editors. Financial Cryptography and Data Security: 11th International Conference, FC 2007, and 1st International Workshop on Usable Security, USEC 2007, Scarborough, Trinidad and Tobago, February 12–16, 2007. Revised Selected Papers, volume 4886 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77365-7 (print), 3-540-77366-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77366-5.

Dicheva:2010:AIM

[DD10] Darina Dicheva and Danail Dochev, editors. Artificial Intelligence: Methodology, Systems, and Applications: 14th International Conference, AIMSA 2010, Varna, Bulgaria, September 8– 10. 2010. Proceedings, volume 6304 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15430-1 (print), 3-642-15431-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15431-7.

Dix:1998:LAI

[DdCF98] J. Dix, Luis Fariñas del Cerro, and Ulrich Furbach, editors. Logics in artificial intelligence: European Workshop, JELIA’98 Dagstuhl, Germany, October 12–15, 1998: Proceedings, volume 1489 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65141-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .E985 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1489.
htm;   http://www.springerlink.
com/content/978-3-540-
65141-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1489.

Dastani:2004:PMA

[DDE04] Mehdi Dastani, Juergen Dix, and Amal El Fallah-Seghrouchni, editors. Programming Multi-Agent Systems: First International Workshop, PROMAS 2003, Melbourne, Australia, July 15, 2003, Selected Revised and Invited Papers, volume 3067 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22180-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3067.htm;
http://www.springerlink.
com/content/978-3-540-
22180-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3067;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98149.

DeGregorio:2005:BVA

[DDFM05] Massimo De Gregorio, Vito Di Maio, Maria Frucci, and Carlo Musio, editors. Brain, vision, and artificial intelligence: first international symposium, BVAI 2005, Naples, Italy, October 19–21, 2005: proceedings, volume 3704 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29282-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP363.3 .B93 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29282-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3704.

Dayde:2005:HPC

[DDHP05] Michel Daydé, Jack J.

Dongarra, Vicente Hernández, and José M. L. M. Palma, editors. High Performance Computing for Computational Science: VECPAR 2004: 6th International Conference Valencia, Spain, June 28–30, 2004 Revised Selected and Invited Papers, volume 3402 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25424-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25424-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3402;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106965.

Dong:2009:ACI

[DDI09] Yingfei Dong, Ding-Zhu Du, and Oscar Ibarra, editors. Algorithms and Computation: 20th International Symposium, ISAAC 2009, Honolulu, Hawaii, USA, December 16–18, 2009. Proceedings, volume 5878 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10630-7 (print), 3-642-10631-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10631-6.

Diamantaras:2010:ANNa

[DDI10a] Konstantinos Diamantaras, Wlodek Duch, and Lazaros S. Iliadis, editors. Artificial Neural Networks — ICANN 2010: 20th International Conference, Thessaloniki, Greece, September 15–18, 2010, Proceedings, Part I, volume 6352 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15818-8 (print), 3-642-15819-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15819-3.

Diamantaras:2010:ANNb

[DDI10b] Konstantinos Diamantaras, Wlodek Duch, and Lazaros S. Iliadis, editors. Artificial Neural Networks — ICANN 2010: 20th International Conference, Thessaloniki, Greece, September 15–18, 2010, Proceedings, Part II, volume 6353 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15821-8 (print), 3-642-15822-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15822-3.

Diamantaras:2010:ANNc

[DDI10c] Konstantinos Diamantaras, Wlodek Duch, and Lazaros S. Iliadis, editors. Artificial Neural Networks — ICANN 2010: 20th International Conference, Thessaloniki, Greece, September 15–18, 2010, Proceedings, Part III, volume 6354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15824-2 (print), 3-642-15825-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15825-4.

Delp:2000:MIC

[DDJ00] Scott L. Delp, Anthony M. DiGoia, and Branislav Jaramaz, editors. Medical image computing and computerassisted intervention — MICCAI 2000: Third International Conference, Pittsburgh, PA, USA, October 11–14, 2000 proceedings, volume 1935 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41189-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R857.O6 I574 2000; QA267.A1 L43 no.1935. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1935.htm;
http://www.springerlink.
com/content/978-3-540-
41189-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1935;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12345.

Dimov:2011:NMA

[DDK11] Ivan Dimov, Stefka Dimova, and Natalia Kolkovska, editors. Numerical Methods and Applications: 7th International Conference, NMA 2010, Borovets, Bulgaria, August 20– 24, 2010. Revised Papers, volume 6046 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18465-0 (print), 3-642-18466-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18466-6.

Degano:2008:CGM

[DDM08] Pierpaolo Degano, Rocco De Nicola, and José Meseguer, editors. Concurrency, Graphs and Models: Essays Dedicated to Ugo Montanari on the Occasion of His 65th Birthday, volume 5065 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68676-2 (print), 3-540-68679-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68679-8.

Dawar:2010:LLI

[DdQ10] Anuj Dawar and Ruy de Queiroz, editors. Logic, Language, Information and Computation: 17th International Workshop, WoL-LIC 2010, Brasilia, Brazil, July 6–9, 2010. Proceedings, volume 6188 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13823-3 (print), 3-642-13824-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13824-9.

Dorigo:2002:AAT

[DDS02] Marco Dorigo, Gianni Di Caro, and Michael Sampels, editors. Ant Algorithms: Third International Workshop, ANTS 2002, Brussels, Belgium, September 12–14, 2002. Proceedings, volume 2463 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44146-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA241 .A43 2002; QA241 .A536 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2463.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44146-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2463.

Danezis:2012:FCDa

[DDS12] George Danezis, Sven Dietrich, and Kazue Sako, editors. Financial Cryptography and Data Security: FC 2011 Workshops, RLCPS and WECSR 2011, Rodney Bay, St. Lucia, February 28 — March 4, 2011, Revised Selected Papers, volume 7126 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29888-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29888-2.

DeGroote:1987:LCB

[De 87] H. F. (Hans F.) De Groote. Lectures on the complexity of bilinear problems, volume 245 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17205-X (paperback), 3-540-17205-X (West Berlin: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). 135 pp. LCCN QA267 .D431 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0245.htm;
 http://www.springerlink.
com/content/978-0-387-
17205-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=245.

DeSantis:1995:ACE

[De 95] Alfredo De Santis, editor. Advances in cryptology — EUROCRYPT ’94: Workshop on the Theory and Application of Cryptographic Techniques, Perugia, Italy, May 9–12, 1994: proceedings, volume 950 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60176-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E965 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0950.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        60176-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=950.

deSwart:1998:ARA

[de 98] Harrie de Swart, editor. Automated reasoning with analytic tableaux and related methods: International Conference, TABLEAUX’98, Oisterwijk, The Netherlands, May 5–8, 1998: Proceedings, volume 1397 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64406-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96T34 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1397.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64406-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1397.

DeMeuter:2007:ASI

[De 07a] Wolfgang De Meuter, editor. Advances in Smalltalk: 14th International Smalltalk Conference, ISC 2006, Prague, Czech Republic, September 4–8, 2006, Revised Selected Papers, volume 4406 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71835-4 (print), 3-540-71836-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71836-9.

DeNicola:2007:PLS

[De 07b] Rocco De Nicola, editor.

Programming Languages and Systems: 16th European Symposium on Programming, ESOP 2007, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2007, Braga, Portugal, March 24 — April 1, 2007. Proceedings, volume 4421 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71314-X (print), 3-540-71316-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71316-6.

DeRaedt:2010:ILP

[De 10a] Luc De Raedt, editor. Inductive Logic Programming: 19th International Conference, ILP 2009, Leuven, Belgium, July 02–04, 2009. Revised Papers, volume 5989 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13839-X (print), 3-642-13840-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13840-9.

DeSchreye:2010:LBP

[De 10b] Danny De Schreye, editor. Logic-Based Program Synthesis and Transformation: 19th International Symposium, LOPSTR 2009, Coimbra, Portugal, September 2009, Revised Selected Papers, volume 6037 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12591-3 (print), 3-642-12592-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-642-
12592-8.

Dearle:2005:CDT

[DE05] Alan Dearle and Susan Eisenbach, editors. Component deployment: Third international working conference, CD 2005, Grenoble, France, November 28–29, 2005, proceedings, volume 3798 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30517-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30517-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3798.

Dechter:2000:PPC

[Dec00] Rina Dechter, editor.

Principles and Practice of Constraint Programming— CP 2000: 6th International Conference, CP 2000, Singapore, September 18–21, 2000. Proceedings, volume 1894 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41053-8 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1894. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1894.htm;
 http://www.springerlink.
com/content/978-3-540-
41053-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1894.

Degano:2003:PLS

[Deg03] Pierpaolo Degano, editor. Programming languages and systems: 12th European Symposium on Programming, ESOP 2003, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2003, Warsaw, Poland, April 7–11, 2003: Proceedings, volume 2618 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00886-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2618.htm;
 http://www.springerlink.
com/content/978-3-540-
00886-6;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2618.

Also available via the World Wide Web.

Degano:2011:FAS

[DEG11] Pierpaolo Degano, Sandro Etalle, and Joshua Guttman, editors. Formal Aspects of Security and Trust: 7th International Workshop, FAST 2010, Pisa, Italy, September 16–17, 2010. Revised Selected Papers, volume 6561 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19750-7 (print), 3-642-19751-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19751-2.

Dehne:1993:ADS

[Deh93] F. Dehne, editor. Algorithms and data structures: third workshop, WADS ’93, Montréal, Canada, August 11– 13, 1993: proceedings, volume 709 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57155-8 (Berlin), 0-387-57155-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 A43 1993. US$86.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0709.
htm;   http://www.springerlink.
com/content/978-3-540-
57155-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=709.

Dehne:1997:ADS

[Deh97] F. (Frank) Dehne, editor. Algorithms and data structures: 5th International Workshop, WADS ’97, Halifax, Nova Scotia, Canada, August 6– 8, 1997: proceedings, volume 1272 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63307-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35W33 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1272.htm;
 http://www.springerlink.
com/content/978-3-540-
63307-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1272.

Dehne:1999:ADS

[Deh99] F. (Frank) Dehne, editor. Algorithms and data structures: 6th International Workshop, WADS’99, Vancouver, Canada, August 11– 14, 1999: Proceedings, volume 1663 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66279-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 W33 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1663.htm;
http://www.springerlink.
com/content/978-3-540-
66279-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1663.

Deimel:1990:SEE

[Dei90] Lionel Deimel, editor.

Software engineering education: SEI Conference 1990, Pittsburgh, Pennsylvania, USA, April 2–3, 1990, proceedings, volume 423 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-97274-9 (New York), 3-540-97274-9 (Berlin) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S6462 1990. DM33.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0423.htm;
 http://www.springerlink.
com/content/978-0-387-
97274-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=423.

“Proceedings of CSEE ’90, the 4th SEI Conference on Software Engineering Education”– Pref.

Dayal:2009:BPM

[DEKR09] Umeshwar Dayal, Johann Eder, Jana Koehler, and Hajo A. Reijers, editors. Business Process Management: 7th International Conference, BPM 2009, Ulm, Germany, September 8– 10, 2009. Proceedings, volume 5701 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03847-6 (print), 3-642-03848-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
03848-8.

DelBimbo:1997:IAPa

[Del97a] Alberto Del Bimbo, editor. Image analysis and processing: 9th international conference, ICIAP ’97, Florence, Italy, September 17–19, 1997: proceedings, volume 1310 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63507-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .I552 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1310.htm;
http://www.springerlink.
com/content/978-3-540-
63507-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1310.

DelBimbo:1997:IAPb

[Del97b] Alberto Del Bimbo, editor. Image analysis and processing: 9th international conference, ICIAP ’97, Florence, Italy, September 17–19, 1997: proceedings, volume 1311 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63508-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .I552 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1311.htm;
 http://www.springerlink.
com/content/978-3-540-
63508-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1311.

Delgado:2000:TIC

[Del00] Jaime Delgado, editor. Telecommunications and IT convergence: towards service evolution: 7th International Conference on Intelligence in Services and Networks: IS and N 2000, Athens, Greece, February 23–25, 2000: Proceedings, volume 1774 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67152-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.A1 I47985 2000; QA267.A1 L43 no.1774.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1774.
htm;   http://www.springerlink.
com/content/978-3-540-
67152-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1774.

Dembinski:1980:MFC

[Dem80] Piotr Dembinski, editor. Mathematical foundations of computer science, 1980: proceedings of the 9th symposium held in Rydzyna, Poland, volume 88 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10027-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S9194 1980. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0088.htm;
http://www.springerlink.
com/content/978-0-387-
10027-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=88.

Dehning:1981:AVM

[DEM81] Waltraud Dehning, Heidrun Essig, and Susanne Maass. The adaptation of virtual mancomputer interfaces to user requirements in dialogs, volume 110 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10826-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 142 pp. LCCN QA76.9.I58 D43. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0110.htm;
 http://www.springerlink.
com/content/978-0-387-
10826-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=110.

“Shortened English version of … Zur Anpassung virtueller Mensch-Rechner-Schnittstellen an Benutzererfordernisse im Dialog”–Pref.

Deza:1996:CCS

[DEM96] M. Deza, R. (Reinhardt) Euler, and I. (Ioannis) Manoussakis, editors. Combinatorics and computer science: 8th Franco-Japanese and 4th Franco-Chinese conference, Brest, France, July 1995: Selected Papers, volume 1120 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61576-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166.C6175 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1120.

htm;   http://www.springerlink.
com/content/978-3-540-
61576-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1120.

Demetrescu:2007:EAI

[Dem07] Camil Demetrescu, editor. Experimental Algorithms: 6th International Workshop, WEA 2007, Rome, Italy, June 6–8, 2007. Proceedings, volume 4525 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72844-9 (print), 3-540-72845-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72845-0.

Denvir:1985:ACS

[Den85] B. T. Denvir, editor. The analysis of concurrent systems: Cambridge, September 12–16, 1983: proceedings, volume 207 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CODEN LNCSD9. ISBN 3-540-16047-7 (West Berlin: paperback), 0-387-16047-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .A4661 1983.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0207.
htm;   http://www.springerlink.
com/content/978-3-540-
16047-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=207.

Deswarte:1992:CSE

[DEQ92] Y. Deswarte, G. Eizenberg, and J.-J. Quisquater, editors. Computer security, ESORICS 92: Second European Symposium on Research in Computer Security, Toulouse, France, November 23–25, 1992: proceedings, volume 648 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-56246-X (Berlin), 0-387-56246-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E973 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0648.htm;
 http://www.springerlink.
com/content/978-3-540-
56246-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=648.

Dershowitz:1989:RTA

[Der89] Nachum Dershowitz, editor. Rewriting techniques and applications: 3rd International Conference, RTA-89, Chapel Hill, North Carolina, USA, April 3–5, 1989: proceedings, volume 355 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51081-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I5551 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0355.htm;
http://www.springerlink.
com/content/978-0-387-
51081-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=355.

Derniame:1992:SPT

[Der92] J. C. Derniame, editor. Software process technology: second European workshop, EWSPT ’92, Trondheim, Norway, September 7–8, 1992: proceedings, volume 635 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55928-0 (Berlin), 0-387-55928-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .E98 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0635.htm;
 http://www.springerlink.
com/content/978-3-540-
55928-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=635.

Driankov:1994:FLF

[DER94] Dimiter Driankov, Peter W. Eklund, and Anca L. Ralescu, editors. Fuzzy logic and fuzzy control: IJCAI ’91 workshops on fuzzy logic and fuzzy control, Sydney, Australia, August 24, 1991: proceedings, volume 833 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58279-7 (Berlin), 0-387-58279-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ212.2 .F89 1994. DM46.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0833.
htm;   http://www.springerlink.
com/content/978-3-540-

58279-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=833.

Dershowitz:2004:VTP

[Der04] Nachum Dershowitz, editor. Verification: Theory and Practice: Essays delivered to Zohar Manna on the occasion of his 64th birthday, volume 2772 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21002-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 V48 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2772.
htm;   http://www.springerlink.
com/content/978-3-540-
21002-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2772;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12001.

Desarmenien:1986:TSD

[Des86] J. Desarmenien, editor. TEX for scientific documentation: second European conference, Strasbourg, France, June 19–21, 1986: proceedings, volume 236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16807-9 (paperback), 3-540-16807-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.4.T47 T491 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0236.htm;
 http://www.springerlink.
com/content/978-0-387-
16807-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=236.

Desmedt:1994:ACC

[Des94] Yvo G. Desmedt, editor. Advances in cryptology, CRYPTO ’94: 14th annual international cryptology conference, Santa Barbara, California, USA, August 21–25, 1994: proceedings, volume 839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58333-5 (Berlin), 0-387-58333-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1994. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t0839.htm;
http://www.springerlink.
com/content/978-3-540-
58333-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=839.

Desmedt:2002:PKC

[Des02] Yvo G. Desmedt, editor.

Public Key Cryptography— PKC 2003: 6th International Workshop on Practice and Theory in Public Key Cryptography Miami, FL, USA, January 6–8, 2003. Proceedings, volume 2567 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00324-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I567 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2567.htm;
http://www.springerlink.
com/content/978-3-540-
00324-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2567.

Also available via the World Wide Web.

Desmedt:2009:ITS

[Des09] Yvo Desmedt, editor. Information Theoretic Security: Second International Conference, ICITS 2007, Madrid, Spain, May 25–29, 2007, Revised Selected Papers, volume 4883 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10229-8 (print), 3-642-10230-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10230-1.

Deussen:1981:TCS

[Deu81] Peter Deussen, editor. Theoretical computer science: 5th GI-conference, Karlsruhe, March 23– 25, 1981, volume 104 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10576-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .T47. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0104.htm;
 http://www.springerlink.
com/content/978-0-387-
10576-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=104.

English, French, or German.

DiazdeIlarrazaSanchez:1996:CAL

[DF96] Arantza Diaz de Ilarraza Sanchez and I. (Isabel) Fernandez de Castro, editors. Computer aided learning and instruction in science and engineering: third international conference, CALISCE ’96, San Sebastian, Spain, July 29– 31, 1996: proceedings, volume 1108 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61491-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q181.A1I538 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1108.
htm;   http://www.springerlink.
com/content/978-3-540-
61491-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1108.

Dzeroski:1999:ILP

[DF99] Saso Dzeroski and Peter A. Flach, editors. Inductive logic programming: 9th International Workshop, ILP-99, Bled, Slovenia, June 24–27, 1999: Proceedings, volume 1634 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66109-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I52 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1634.htm;
 http://www.springerlink.
com/content/978-3-540-
66109-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1634.

Danvy:2001:PDO

[DF01a] Olivier Danvy and Andrzej Filinski, editors. Programs as data objects: Second Symposium, PADO 2001, Aarhus, Denmark, May 21–23, 2001: Proceedings, volume 2053 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42068-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .P33 2001; QA267.A1 L43 no.2053. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t2053.htm;
http://www.springerlink.
com/content/978-3-540-
42068-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2053.

Davida:2001:ISI

[DF01b] George Davida and Yair Frankel, editors. Information security: 4th International Conference, ISC 2001, Malaga, Spain, October 1–3, 2001: Proceedings, volume 2200 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42662-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2200.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2200.
htm;   http://www.springerlink.
com/content/978-3-540-
42662-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2200.

Outgrowth of the Information Security Workshop (ISW) series.

Domingo-Ferrer:2002:ICS

[DF02] Josep Domingo-Ferrer.

Inference Control in Statistical Databases: From Theory to Practice, volume 2316 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43614-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 229 pp. LCCN QA76.9.D314 I54 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2316.
htm;   http://www.springerlink.
com/content/978-3-540-
43614-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2316.

Dourish:2006:UUC

[DF06] Paul Dourish and Adrian Friday, editors. UbiComp 2006: Ubiquitous Computing: 8th International Conference, UbiComp 2006 Orange County, CA, USA, September 17– 21, 2006 Proceedings, volume 4206 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39634-9 (print), 3-540-39635-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
39635-2.

DeLucia:2009:SEI

[DF09] Andrea De Lucia and Filomena Ferrucci, editors. Software Engineering: International Summer Schools, ISSSE 2006-2008, Salerno, Italy, Revised Tutorial Lectures, volume 5413 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-95887-8 (print), 3-540-95888-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
95888-8.

Daniel:2010:CTW

[DF10] Florian Daniel and Federico Michele Facca, editors. Current Trends in Web Engineering: 10th International Conference on Web Engineering ICWE 2010 Workshops, Vienna, Austria, July 2010, Revised Selected Papers, volume 6385 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16984-8 (print), 3-642-16985-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16985-4.

Delgrande:2011:LPN

[DF11] James P. Delgrande and Wolfgang Faber, editors. Logic Programming and Nonmonotonic Reasoning: 11th International Conference, LPNMR 2011, Vancouver, Canada, May 16– 19, 2011. Proceedings, volume 6645 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20894-0 (print), 3-642-20895-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20895-9.

Downey:2004:PEC

[DFD04] Rod Downey, Michael Fellows, and Frank Dehne, editors. Parameterized and Exact Computation: First International Workshop, IW-PEC 2004, Bergen, Norway, September 14–17, 2004, Proceedings, volume 3162 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23071-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.8 .I97 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3162.htm;
 http://www.springerlink.
com/content/978-3-540-
23071-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3162;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100584.

deFrutos-Escrig:2004:FTN

[dFEN04] David de Frutos-Escrig and Manuel Núñez, editors. Formal Techniques for Networked and Distributed Systems— FORTE 2004: 24th IFIP WG 6.1 International Conference Madrid, Spain, September 27–30, 2004 Proceedings, volume 3235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23252-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 F67 2004.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t3235.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        23252-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3235;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100576.

Domingo-Ferrer:2006:PSD

[DFF06] Josep Domingo-Ferrer and Luisa Franconi, editors. Privacy in Statistical Databases: CENEX-SDC Project International Conference, PSD 2006, Rome, Italy, December 13–15, 2006. Proceedings, volume 4302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49330-1 (print), 3-540-49332-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49332-7.

Derrick:2012:ASM

[DFG^{+}12] John Derrick, John Fitzgerald, Stefania Gnesi, Sarfraz Khurshid, Michael Leuschel, Steve Reeves, and Elvinia Riccobene, editors. Abstract State Machines, Alloy, B, VDM, and Z: Third International Conference, ABZ 2012, Pisa, Italy, June 18–21, 2012. Proceedings, volume 7316 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30884-8 (print), 3-642-30885-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30885-7.

Durrleman:2012:STI

[DFGN12] Stanley Durrleman, Tom Fletcher, Guido Gerig, and Marc Niethammer, editors. Spatio-temporal Image Analysis for Longitudinal and Time-Series Image Data: Second International Workshop, STIA 2012, Held in Conjunction with MIC-CAI 2012, Nice, France, October 1, 2012. Proceedings, volume 7570 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33554-3 (print), 3-642-33555-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        33555-6.

Dehne:1991:ACI

[DFK91] F. Dehne, F. Fiala, and Waldemar W. Koczkodaj, editors. Advances in computing and information — ICCI ’91: International Conference on Computing and Information, Ottawa, Canada, May 27–29, 1991: proceedings, volume 497 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54029-6 (New York), 3-540-54029-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I587 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0497.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        54029-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=497.

DeRaedt:2008:PIL

[DFKM08] Luc De Raedt, Paolo Frasconi, Kristian Kersting, and Stephen Muggleton, editors. Probabilistic Inductive Logic Programming: Theory and Applications, volume 4911 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78651-1 (print), 3-540-78652-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78652-8.

Daehlen:2010:MMC

[DFL^{+}10] Morten Dæhlen, Michael Floater, Tom Lyche, Jean-Louis Merrien, Knut Mørken, and Larry L. Schumaker, editors. Mathematical Methods for Curves and Surfaces: 7th International Conference, MMCS 2008, Tønsberg, Norway, June 26–July 1, 2008, Revised Selected Papers, volume 5862 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11619-1 (print), 3-642-11620-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11620-9.

DeNicola:2004:CML

[DFM04] Rocco De Nicola, Gianluigi Ferrari, and Greg Meredith, editors. Coordination Models and Languages: 6th International Conference, CO-ORDINATION 2004, Pisa, Italy, February 24–27, 2004: Proceedings, volume 2949 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21044-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .C695 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2949.htm;
 http://www.springerlink.
com/content/978-3-540-
21044-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2949;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95570.

Domingo-Ferrer:2011:PSD

[DFM11] Josep Domingo-Ferrer and Emmanouil Magkos, editors. Privacy in Statistical Databases: UN-ESCO Chair in Data Privacy, International Conference, PSD 2010, Corfu, Greece, September 22–24, 2010. Proceedings, volume 6344 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-15837-4 (print), 3-642-15838-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15838-4.

Dediu:2010:LAT

[DFMV10] Adrian-Horia Dediu, Henning Fernau, and Carlos Martín-Vide, editors. Language and Automata Theory and Applications: 4th International Conference, LATA 2010, Trier, Germany, May 24–28, 2010. Proceedings, volume 6031 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13088-7 (print), 3-642-13089-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13089-2.

Dix:1997:LPN

[DFN97] J. (Jurgen) Dix, Ulrich Fuhrbach, and Anil Nerode, editors. Logic programming and nonmonotonic reasoning:

4th International Conference, LPNMR ’97, Dagstuhl Castle, Germany, July 1997: proceedings, volume 1265 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63255-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63.L74 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1265.
htm;   http://www.springerlink.
com/content/978-3-540-
63255-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1265.

Dix:2010:CLMa

[DFN10] Jürgen Dix, Michael Fisher, and Peter Novák, editors. Computational Logic in Multi-Agent Systems: 10th International Workshop, CLIMA X, Hamburg, Germany, September 9– 10, 2009, Revised Selected and Invited Papers, volume 6214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16866-3 (print), 3-642-16867-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16867-3.

Dellaert:2012:OLS

[DFP^{+}12] Frank Dellaert, Jan-Michael Frahm, Marc Pollefeys, Laura Leal-Taixé, and Bodo Rosenhahn, editors. Outdoor and Large-Scale Real-World Scene Analysis: 15th International Workshop on Theoretical Foundations of Computer Vision, Dagstuhl Castle, Germany, June 26 — July 1, 2011. Revised Selected Papers, volume 7474 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34090-3 (print), 3-642-34091-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34091-8.

Domingo-Ferrer:2006:SCR

[DFPS06] Josep Domingo-Ferrer, Joachim Posegga, and Daniel Schreckling, editors. Smart Card Research and Advanced Applications: 7th IFIP WG 8.8/11.2 International Conference, CARDIS 2006, Tarragona, Spain, April 19–21, 2006. Proceedings, volume 3928 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33311-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33311-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3928.

DeVos:2011:COI

[DFPV11] Marina De Vos, Nicoletta Fornara, Jeremy V. Pitt, and George Vouros, editors. Coordination, Organizations, Institutions, and Norms in Agent Systems VI: COIN 2010 International Workshops, COIN@AAMAS 2010, Toronto, Canada, May 2010, COIN@MALLOW 2010, Lyon, France, August 2010, Revised Selected Papers, volume 6541 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21267-0 (print), 3-642-21268-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21268-0.

Davida:2002:ISI

[DFR02] George Davida, Yair Frankel, and Owen Rees, editors. Infrastructure Security: International Conference, InfraSec 2002 Bristol, UK, October 1–3, 2002. Proceedings, volume 2437 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44309-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I552 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2437.htm;
http://www.springerlink.
com/content/978-3-540-
44309-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2437.

Dustdar:2006:BPM

[DFS06] Schahram Dustdar, José Luiz Fiadeiro, and Amit P. Sheth, editors. Business Process Management: 4th International Conference, BPM 2006, Vienna, Austria, September 5–7, 2006. Proceedings, volume 4102 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38901-6 (print), 3-540-38903-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38903-3.

Domingo-Ferrer:2008:PSD

[DFS08] Josep Domingo-Ferrer and Yücel Saygın, editors. Privacy in Statistical Databases: UNESCO Chair in Data Privacy International Conference, PSD 2008, Istanbul, Turkey, September 24–26, 2008. Proceedings, volume 5262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87470-4 (print), 3-540-87471-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87471-3.

Domingo-Ferrer:2004:PSD

[DFT04] Josep Domingo-Ferrer and Vicenç Torra, editors. Privacy in Statistical Databases: CASC Project International Workshop, PSD 2004, Barcelona, Spain, June 9–11, 2004, Proceedings, volume 3050 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22118-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3050.htm;
 http://www.springerlink.
com/content/978-3-540-
22118-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3050;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97945.

Domingue:2009:FIF

[DFT09] John Domingue, Dieter Fensel, and Paolo Traverso, editors. Future Internet — FIS 2008: First Future Internet Symposium, FIS 2008 Vienna, Austria, September 29–30, 2008 Revised Selected Papers, volume 5468 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00984-0 (print), 3-642-00985-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00985-3.

Domingo-Ferrer:2012:PSD

[DFT12] Josep Domingo-Ferrer and Ilenia Tinnirello, editors. Privacy in Statistical Databases: UN-ESCO Chair in Data Privacy, International Conference, PSD 2012, Palermo, Italy, September 26–28, 2012. Proceedings, volume 7556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33626-4 (print), 3-642-33627-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33627-0.

Davenport:2011:ICM

[DFUR11] James H. Davenport, William M. Farmer, Josef Urban, and Florian Rabe, editors. Intelligent Computer Mathematics: 18th Symposium, Calculemus 2011, and 10th International Conference, MKM 2011, Bertinoro, Italy, July 18–23, 2011. Proceedings, volume 6824 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22672-8 (print), 3-642-22673-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22673-1.

Du:2011:WTA

[DFW^{+}11] Xiaoyong Du, Wenfei Fan, Jianmin Wang, Zhiyong Peng, and Mohamed A. Sharaf, editors. Web Technologies and Applications: 13th Asia-Pacific Web Conference, APWeb 2011, Beijing, China, April 18–20, 2011. Proceedings, volume 6612 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20290-X (print), 3-642-20291-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        20291-9.

Dawson:1996:CPA

[DG96] Ed (Edward) Dawson and Jovan Golic, editors. Cryptography: policy and algorithms: international conference, Brisbane, Queensland, Australia, July 3–5, 1995: proceedings, volume 1029 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60759-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C844 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1029.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60759-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1029.

Duncan:1997:IPM

[DG97] James Duncan and Gene Gindi, editors. Information processing in medical imaging: 15th International Conference, IPMI’97, Poultney, Vermont, USA, June 9– 13, 1997, proceedings, volume 1230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63046-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN RC78.7.D53 I573 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1230.htm;
 http://www.springerlink.
com/content/978-3-540-
63046-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1230.

Dignum:2000:IAC

[DG00] Frank Dignum and Mark Greaves. Issues in agent communication, volume 1916 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41144-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 350 pp. LCCN QA76.76.I58 I595 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1916.
htm;   http://www.springerlink.
com/content/978-3-540-
41144-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1916.

DeAlfaro:2001:PAP

[DG01] Luca De Alfaro and Stephen Gilmore, editors. Process algebra and probabilistic methods: performance modeling and verification: joint International Workshop, PAPM-PROBMIV 2001, Aachen, Germany, September 12–14, 2001: Proceedings, volume 2165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42556-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2165; QA267 .P36 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2165.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42556-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2165.

Dean:2002:CBS

[DG02] John Dean and Andree Gravel, editors. COTS-Based Software Systems: First International Conference, IC-CBSS 2002, Orlando, FL, USA, February 4– 6, 2002: Proceedings, volume 2255 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43100-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .I33 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2255.htm;
 http://www.springerlink.
com/content/978-3-540-
43100-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2255.

Das:2004:IIT

[DG04] Gautam Das and V. P.

(Ved Prakash) Gulati, editors. Intelligent information technology: 7th International Conference on Information Technology, CIT 2004, Hyderabad, India, December 20–23, 2004: Proceedings, volume 3356 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24126-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T58.5 .I58 2004. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3356;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        24126-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3356;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104585.

Duncan:2005:MICb

[DG05a] James S. Duncan and Guido Gerig, editors. Medical Image Computing and Computer-Assisted Intervention

— MICCAI 2005, volume 3750 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29326-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29326-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3750.

Duncan:2005:MICa

[DG05b] James S. Duncan and Guido Gerig, editors. Medical image computing and computer-assisted intervention — MIC-CAI 2005: 8th international conference, Palm Springs, CA, USA, October 26–29, 2005, part I, volume 3749 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29327-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29327-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3749.

Danezis:2006:PETb

[DG06a] George Danezis and Philippe Golle, editors. Privacy Enhancing Technologies: 6th International Workshop, PET 2006, Cambridge, UK, June 28–30, 2006, Revised Selected Papers, volume 4258 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68790-4 (print), 3-540-68793-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68793-1.

Datta:2006:SSS

[DG06b] Ajoy K. Datta and Maria Gradinariu, editors. Stabilization, Safety, and Security of Distributed Systems: 8th International Symposium, SSS 2006, Dallas, TX, USA, November 17–19, 2006. Proceedings, volume 4280 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49018-3 (print), 3-540-49823-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49823-0.

Davies:2007:IFM

[DG07a] Jim Davies and Jeremy Gibbons, editors. Integrated Formal Methods: 6th International Conference, IFM 2007, Oxford, UK, July 2–5, 2007. Proceedings, volume 4591 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73209-8 (print), 3-540-73210-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73210-5.

Deng:2007:INE

[DG07b] Xiaotie Deng and Fan Chung Graham, editors. Internet and Network Economics: Third International Workshop, WINE 2007, San Diego, CA, USA, December 12–14, 2007. Proceedings, volume 4858 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77104-2 (print), 3-540-77105-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77105-0.

Degano:2009:CMS

[DG09a] Pierpaolo Degano and Roberto Gorrieri, editors. Computational Methods in Systems Biology: 7th International Conference, CMSB 2009, Bologna, Italy, August 31–September 1, 2009. Proceedings, volume 5688 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03844-1 (print), 3-642-03845-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03845-7.

Dingledine:2009:FCD

[DG09b] Roger Dingledine and Philippe Golle, editors. Financial Cryptography and Data Security: 13th International Conference, FC 2009, Accra Beach, Barbados, February 23–26, 2009. Revised Selected Papers, volume 5628 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03548-5 (print), 3-642-03549-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03549-4.

Degano:2010:FAS

[DG10] Pierpaolo Degano and Joshua D. Guttman, editors. Formal Aspects in Security and Trust: 6th International Workshop, FAST 2009, Eindhoven, The Netherlands, November 5–6, 2009, Revised Selected Papers, volume 5983 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12458-5 (print), 3-642-12459-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12459-4.

Degano:2012:PST

[DG12] Pierpaolo Degano and Joshua D. Guttman, editors. Principles of Security and Trust: First International Conference, POST 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012, Proceedings, volume 7215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28640-2 (print), 3-642-28641-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28641-4.

Dorigo:2006:ACO

[DGB^{+}06] Marco Dorigo, Luca Maria Gambardella, Mauro Birattari, Alcherio Martinoli, Riccardo Poli, and Thomas Stützle, editors. Ant Colony Optimization and Swarm Intelligence: 5th International Workshop, ANTS 2006, Brussels, Belgium, September 4–7, 2006. Proceedings, volume 4150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38482-0 (print), 3-540-38483-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38483-0.

Dimitrakakis:2011:PSI

[DGDM^{+}11] Christos Dimitrakakis, Aris Gkoulalas-Divanis, Aikaterini Mitrokotsa, Vassilios S. Verykios, and Yücel Saygin, editors. Privacy and Security Issues in Data Mining and Machine Learning: International ECML/ PKDD Workshop, PS-DML 2010, Barcelona, Spain, September 24, 2010. Revised Selected Papers, volume 6549 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19895-3 (print), 3-642-19896-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19896-0.

deGroote:2011:FGI

[dGEK11] Philippe de Groote, Markus Egg, and Laura Kallmeyer, editors. Formal Grammar: 14th International Conference, FG 2009, Bordeaux, France, July 25–26, 2009, Revised Selected Papers, volume 5591 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20168-7 (print), 3-642-20169-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20169-1.

Doncieux:2010:AAI

[DGG^{+}10] Stéphane Doncieux, Benoît Girard, Agnès Guillot, John Hallam, Jean-Arcady Meyer, and Jean-Baptiste Mouret, editors. From Animals to Animats 11: 11th International Conference on Simulation of Adaptive Behavior, SAB 2010, Paris — Clos Lucé, France, August 25–28, 2010. Proceedings, volume 6226 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15192-2 (print), 3-642-15193-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15193-4.

Declerck:2011:SMI

[DGG^{+}11a] Thierry Declerck, Michael Granitzer, Marcin Grzegorzek, Massimo Romanelli, Stefan Rüger, and Michael Sintek, editors. Semantic Multimedia: 5th International Conference on Semantic and Digital Media Technologies, SAMT 2010. Saarbrücken, Germany, December 1–3, 2010. Revised Selected Papers, volume 6725 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23016-4 (print), 3-642-23017-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23017-2.

Domingue:2011:FIF

[DGG^{+}11b] John Domingue, Alex Galis, Anastasius Gavras, Theodore Zahariadis, Dave Lambert, Frances Cleary, Petros Daras, Srdjan Krco, Henning Müller, Man-Sze Li, Hans Schaffers, Volkmar Lotz, Federico Alvarez, Burkhard Stiller, Stamatis Karnouskos, Susanna Avessta, and Michael Nilsson, editors. The Future Internet: Future Internet Assembly 2011: Achievements and Technological Promises, volume 6656 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20897-5 (print), 3-642-20898-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20898-0.

Dean:2005:WIS

[DGJ^{+}05] Mike Dean, Yuanbo Guo, Woochun Jun, Roland Kaschek, Shonali Krishnaswamy, Zhengxiang Pan, and Quan Z. Sheng, editors. Web Information Systems Engineering — WISE 2005 Workshops: WISE 2005 International Workshops, New York, NY, USA, November 20–22, 2005. Proceedings, volume 3807 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30018-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30018-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3807.

Dou:2009:APP

[DGJ09] Yong Dou, Ralf Gruber, and Josef M. Joller, editors. Advanced Parallel Processing Technologies: 8th International Symposium, APPT 2009, Rapperswil, Switzerland, August 24–25, 2009 Proceedings, volume 5737 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03643-0 (print), 3-642-03644-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03644-6.

Derrick:2012:IFM

[DGLT12] John Derrick, Stefania Gnesi, Diego Latella, and Helen Treharne, editors. Integrated Formal Methods: 9th International Conference, IFM 2012, Pisa, Italy, June 18–21, 2012. Proceedings, volume 7321 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30728-0 (print), 3-642-30729-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30729-4.

Daelemans:2008:MLKa

[DGM08a] Walter Daelemans, Bart Goethals, and Katharina Morik, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2008, Antwerp, Belgium, September 15– 19, 2008, Proceedings, Part I, volume 5211 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87478-X (print), 3-540-87479-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87479-9.

Daelemans:2008:MLKb

[DGM08b] Walter Daelemans, Bart Goethals, and Katharina Morik, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2008, Antwerp, Belgium, September 15– 19, 2008, Proceedings, Part II, volume 5212 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87480-1 (print), 3-540-87481-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87481-2.

Degano:2009:FASa

[DGM09] Pierpaolo Degano, Joshua Guttman, and Fabio Martinelli, editors. Formal Aspects in Security and Trust: 5th International Workshop, FAST 2008 Malaga, Spain, October 9–10, 2008 Revised Selected Papers, volume 5491 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01464-X (print), 3-642-01465-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01465-9.

Degano:1997:ALP

[DGMS97] Pierpaolo Degano, Roberto Gorrieri, and Alberto Marchetti-Spaccamela, editors. Automata, languages, and programming: 24th international colloquium, ICALP’97, Bologna, Italy, July 7– 11, 1997: proceedings, volume 1256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63165-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I23 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1256.htm;
 http://www.springerlink.
com/content/978-3-540-
63165-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1256.

The International Colloquium on Algorithms, Languages, and Programming (ICALP) is the annual conference of the European Association for Theoretical Computer Science (EATCS).

Dittrich:2001:AIS

[DGN01] Klaus R. Dittrich, Andreas Geppert, and Moira C. Norrie, editors. Advanced information systems engineering: 13th International Conference, CAiSE 2001, Interlaken, Switzerland, June 4–8, 2001: Proceedings, volume 2068 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42215-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .C33 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2068.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42215-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2068.

deGroote:2012:FGI

[dGN12] Philippe de Groote and Mark-Jan Nederhof, editors. Formal Grammar: 15th and 16th International Conferences, FG 2010, Copenhagen, Denmark, August 2010, FG 2011, Ljubljana, Slovenia, August 2011, Revised Selected Papers, volume 7395 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32023-6 (print), 3-642-32024-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32024-8.

Dolev:1992:TCS

[DGR92] D. Dolev, Zvi Galil, and M. Rodeh, editors. Theory of computing and systems: ISTCS ’92, Israel symposium, Haifa, Israel, May 27–28, 1992: proceedings, volume 601 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55553-6 (Berlin), 0-387-55553-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I76 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0601.
htm;   http://www.springerlink.
com/content/978-3-540-

                                        55553-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=601.

Dikenelli:2006:ESA

[DGR06] Oguz Dikenelli, Marie-Pierre Gleizes, and Alessandro Ricci, editors. Engineering Societies in the Agents World VI: 6th International Workshop, ESAW 2005, Kusadasi, Turkey, October 26–28, 2005, Revised Selected and Invited Papers, volume 3963 of Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34451-9 (soft cover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34451-3.

DMello:2011:ACIa

[DGSM11a] Sidney D’Mello, Arthur Graesser, Björn Schuller, and Jean-Claude Martin, editors. Affective Computing and Intelligent Interaction: 4th International Conference, ACII 2011, Memphis, TN, USA, October 9– 12, 2011, Proceedings, Part I, volume 6974 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24599-4 (print), 3-642-24600-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24600-5.

DMello:2011:ACIb

[DGSM11b] Sidney D’Mello, Arthur Graesser, Björn Schuller, and Jean-Claude Martin, editors. Affective Computing and Intelligent Interaction: Fourth International Conference, ACII 2011, Memphis, TN, USA, October 9– 12, 2011, Proceedings, Part II, volume 6975 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24570-6 (print), 3-642-24571-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24571-8.

Detyniecki:2011:AMRa

[DGSN11] Marcin Detyniecki, Ana García-Serrano, and Andreas Nürnberger, editors. Adaptive Multimedia Retrieval. Understanding Media and Adapting to the User: 7th International Workshop, AMR 2009, Madrid, Spain, September 24–25, 2009, Revised Selected Papers, volume 6535 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18448-0 (print), 3-642-18449-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18449-9.

Dehne:2009:ADS

[DGST09] Frank Dehne, Marina Gavrilova, Jörg-Rüdiger Sack, and Csaba D. Tóth, editors. Algorithms and Data Structures: 11th International Symposium, WADS 2009, Banff, Canada, August 21–23, 2009. Proceedings, volume 5664 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03366-0 (print), 3-642-03367-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03367-4.

Danvy:1996:PEI

[DGT96] Olivier Danvy, R. (Robert)

Gluck, and P. (Peter) Thiemann, editors. Partial Evaluation International Seminar, Dagstuhl Castle, Germany, February 1996: Selected Papers, volume 1110 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61580-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E93 P37 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1110.
htm;   http://www.springerlink.
com/content/978-3-540-
61580-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1110.

DAmbra:2010:EPPa

[DGT10a] Pasqua D’Ambra, Mario Guarracino, and Domenico Talia, editors. Euro-Par 2010 — Parallel Processing: 16th International Euro-Par Conference, Ischia, Italia, August 31 — September 3, 2010, Proceedings, Part I, volume 6271 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15276-7 (print), 3-642-15277-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15277-1.

DAmbra:2010:EPPb

[DGT10b] Pasqua D’Ambra, Mario Guarracino, and Domenico Talia, editors. Euro-Par 2010 — Parallel Processing: 16th International Euro-Par Conference, Ischia, Italy, August 31 — September 3, 2010, Proceedings, Part II, volume 6272 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15290-2 (print), 3-642-15291-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15291-7.

Dan:2010:SOC

[DGT10c] Asit Dan, Frédéric Gittler, and Farouk Toumani, editors. Service-Oriented Computing. ICSOC/ServiceWave 2009 Workshops: International Workshops, IC-SOC/ServiceWave 2009, Stockholm, Sweden, November 23–27, 2009, Revised Selected Papers, volume 6275 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16131-6 (print), 3-642-16132-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16132-2.

Dias:2009:GBH

[DGWB09] Miguel Sales Dias, Sylvie Gibet, Marcelo M. Wanderley, and Rafael Bastos, editors. Gesture-Based Human-Computer Interaction and Simulation: 7th International Gesture Workshop, GW 2007, Lisbon, Portugal, May 23–25, 2007, Revised Selected Papers, volume 5085 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-92864-2 (print), 3-540-92865-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92865-2.

Diaz-Herrera:1994:SEE

[DH94] Jorge L. Diaz-Herrera, editor. Software engineering education: 7th SEI CSEE Conference, San Antonio, Texas, USA, January 5–7, 1994: proceedings, volume 750 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57461-1 (Berlin), 0-387-57461-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S45 1994. DM114.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0750.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        57461-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=750.

“7th Conference on Software Engineering Education”–Pref.

DeGroote:1997:TLC

[DH97] Philippe De Groote and J. Roger Hindley, editors. Typed lambda calculi and applications: Third International Conference on Typed Lambda Calculi and Applications, TLCA ’97, Nancy, France, April 2–4, 1997: proceedings, volume 1210 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62688-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.5 .I58 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1210.
htm;   http://www.springerlink.
com/content/978-3-540-
62688-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1210.

Datta:2001:SSS

[DH01] Ajoy K. Datta and Ted Herman, editors. Selfstabilizing systems: 5th International Workshop, WSS 2001, Lisbon, Portugal, October 1–2, 2001: Proceedings, volume 2194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42653-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2194. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2194.htm;
 http://www.springerlink.
com/content/978-3-540-
42653-0;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2194.

Diekert:2004:SAS

[DH04] Volker Diekert and Michel Habib, editors. STACS 2004: 21st Annual Symposium on Theoretical Aspects of Computer Science, Montpellier, France, March 25– 27, 2004: Proceedings, volume 2996 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21236-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S956 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2996.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21236-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2996;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96012.

DalmauRoyo:2005:MMN

[DH05] Jordi Dalmau Royo and Go Hasegawa, editors. Management of multimedia networks and services: 8th International Conference on Management of Multimedia Networks and Services, MMNS 2005, Barcelona, Spain, October 24–26, 2005: proceedings, volume 3754 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29641-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.75 .I58 2005; TK5105.15 .I58 2005.

URL http://www.springerlink.
com/content/978-3-540-
29641-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3754.

Dix:2006:FIK

[DH06] Jürgen Dix and Stephen J.

Hegner, editors. Foundations of information and knowledge systems: 4th international symposium, FoIKS 2006, Budapest, Hungary, February 14–17, 2006: proceedings, volume 3861 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31782-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .F65 2006. URL http:

//www.springerlink.
com/content/978-3-540-

                                        31782-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3861.

Damm:2007:CAV

[DH07a] Werner Damm and Holger Hermanns, editors. Computer Aided Verification: 19th International Conference, CAV 2007, Berlin, Germany, July 3–7, 2007. Proceedings, volume 4590 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73367-1 (print), 3-540-73368-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73368-3.

Duparc:2007:CSL

[DH07b] Jacques Duparc and Thomas A. Henzinger, editors. Computer Science Logic: 21st International Workshop, CSL 2007, 16th Annual Conference of the EACSL, Lausanne, Switzerland, September 11–15, 2007. Proceedings, volume 4646 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74914-4 (print), 3-540-74915-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74915-8.

Dumas:2008:WSF

[DH08] Marlon Dumas and Reiko Heckel, editors. Web Services and Formal Methods: 4th International Workshop, WS-FM 2007, Brisbane, Australia, September 28– 29, 2007. Proceedings, volume 4937 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79229-5 (print), 3-540-79230-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79230-7.

Deshpande:2010:SUM

[DH10a] Amol Deshpande and Anthony Hunter, editors. Scalable Uncertainty Management: 4th International Conference, SUM 2010, Toulouse, France, September 27–29, 2010. Proceedings, volume 6379 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15950-8 (print), 3-642-15951-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15951-0.

DHondt:2010:EOO

[D’H10b] Theo D’Hondt, editor. ECOOP 2010 — Object-Oriented Programming: 24th European Conference, Maribor, Slovenia, June 21–25, 2010. Proceedings, volume 6183 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14106-4 (print), 3-642-14107-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14107-2.

Demetrescu:2011:AEA

[DH11] Camil Demetrescu and Magnús M. Halldórsson, editors. Algorithms — ESA 2011: 19th Annual European Symposium, Saarbrücken, Germany, September 5–9, 2011. Proceedings, volume 6942 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23718-5 (print), 3-642-23719-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23719-5.

Duke:2008:SMT

[DHH^{+}08] David Duke, Lynda Hardman, Alex Hauptmann, Dietrich Paulus, and Steffen Staab, editors. Semantic Multimedia: Third International Conference on Semantic and Digital Media Technologies, SAMT 2008, Koblenz, Germany, December 3–5, 2008. Proceedings, volume 5392 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92234-2 (print), 3-540-92235-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92235-3.

Duke:1999:PFM

[DHM99] David J. Duke, I. Herman, and Scott Marshall.

PREMO: A Framework for Multimedia Middleware: Specification, Rationale, and Java Binding, volume 1591 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66720-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 254 pp. LCCN QA76.575 .D85 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1591.
htm;   http://www.springerlink.
com/content/978-3-540-
66720-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1591.

Deransart:2000:AVT

[DHM00] Pierre Deransart, Manuel Hermenegildo, and J. Małuszynski. Analysis and visualization tools for constraint programming: constraint debugging, volume 1870 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41137-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xxi + 363 pp. LCCN QA76.612 .A53 2000.

URL http://link.springer-
ny.com/link/service/

series/0558/tocs/t1870.
htm;   http://www.springerlink.
com/content/978-3-540-
41137-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1870.

Dowek:1996:HOA

[DHMM96] Gilles Dowek, Jan Heering, Karl Meinke, and Bernhard Möller, editors. Higher-Order Algebra, Logic, and Term Rewriting: Second International Workshop, HOA ’95, Paderborn, Germany, September 21– 22, 1995: Selected Papers, volume 1074 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61254-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .H484 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1074.htm;
http://www.springerlink.
com/content/978-3-540-
61254-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1074.

Dolev:2008:OSF

[DHO08] Shlomi Dolev, Tobias Haist, and Mihai Oltean, editors. Optical Super-Computing: First International Workshop, OSC 2008, Vienna, Austria, August 26, 2008. Proceedings, volume 5172 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85672-2 (print), 3-540-85673-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85673-3.

Du:2009:COA

[DHP09] Ding-Zhu Du, Xiaodong Hu, and Panos M. Pardalos, editors. Combinatorial Optimization and Applications: Third International Conference, COCOA 2009, Huangshan, China, June 10– 12, 2009. Proceedings, volume 5573 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02025-9 (print), 3-642-02026-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02026-1.

Dean:2010:SWR

[DHRT10] Mike Dean, John Hall, Antonino Rotolo, and Said Tabet, editors. Semantic Web Rules: International Symposium, RuleML 2010, Washington, DC, USA, October 21–23, 2010. Proceedings, volume 6403 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16288-6 (print), 3-642-16289-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16289-3.

Dams:2010:CCC

[DHS10] Dennis Dams, Ulrich Hannemann, and Martin Steffen, editors. Concurrency, Compositionality, and Correctness: Essays in Honor of Willem-Paul de Roever, volume 5930 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11511-X (print), 3-642-11512-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        11512-7.

Dyckhoff:1996:ELP

[DHSH96] Roy Dyckhoff, Heinrich Herre, and Peter Joseph Schroeder-Heister, editors. Extensions of logic programming: 5th international workshop, ELP ’96, Leipzig, Germany, March 28–30, 1996: proceedings, volume 1050 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60983-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .E47 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1050.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60983-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1050.

Dechesne:2012:AAT

[DHtM^{+}12] Francien Dechesne, Hiromitsu Hattori, Adriaan ter Mors, Jose Miguel Such, Danny Weyns, and Frank Dignum, editors. Advanced Agent Technology: AAMAS 2011 Workshops, AM-PLE, AOSE, ARMS, DOCM 3 AS, ITMAS, Taipei, Taiwan, May 2– 6, 2011. Revised Selected Papers, volume 7068 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27215-0 (print), 3-642-27216-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27216-5.

Deng:2009:FAT

[DHX09] Xiaotie Deng, John E.

Hopcroft, and Jinyun Xue, editors. Frontiers in algorithmics. Third international workshop, FAW 2009, Hefei, China, June 20–23, 2009. Proceedings, volume 5598 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-642-02270-7, 3-642-02269-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 F39 2009.

URL http://www.springerlink.
com/content/978-3-642-
02270-8;   http://www.
springerlink.com/content/
p1040n9jt618;   http:
//www.zentralblatt-
math.org/zmath/en/search/
?an=1166.68003.

DiBattista:1997:GDI

[Di 97] Giuseppe Di Battista, editor. Graph drawing: 5th International Symposium, GD’97, Rome, Italy, September 18– 20, 1997: proceedings, volume 1353 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63938-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1353.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1353.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63938-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1353.

Diaz:1983:ALP

[Dia83] J. Diaz, editor. Automata, languages, and programming: 10th colloquium, Barcelona, Spain, July 18–22, 1983, volume 154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-12317-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.154. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0154.htm;
 http://www.springerlink.
com/content/978-0-387-
12317-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=154.

Proceedings of the Tenth International Colloquium on Automata, Languages, and Programming, sponsored by the European Association for Theoretical Computer Science.

Diekert:1990:CT

[Die90] Volker Diekert. Combinatorics on traces, volume 454 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-53031-2, 0-387-53031-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 164 pp. LCCN QA267.5.S4 D54 1990. DM33.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0454.htm;
 http://www.springerlink.
com/content/978-3-540-
53031-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=454.

Diehl:2002:SVI

[Die02] Stephan Diehl, editor. Software visualization: international seminar, Dagstuhl Castle, Germany, May 20– 25, 2001: Revised Papers, volume 2269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43323-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .D34 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2269.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43323-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2269.

Dietz:2003:LCP

[Die03] Henry G. Dietz, editor. Languages and Compilers for Parallel Computing: 14th International Workshop, LCPC 2001, Cumberland Falls, KY, USA, August 1– 3, 2001: Revised Papers, volume 2624 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-04029-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2624.htm;
 http://www.springerlink.
com/content/978-3-540-
04029-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2624.

The 14th workshop on Languages and Compilers for Parallel Computing, LCPC 2001, was organized and hosted by the Electrical and Computer Engineering Department of the University of Kentucky, Lexington, KY, USA.

Dietzfelbinger:2004:PTP

[Die04] Martin Dietzfelbinger.

Primality Testing in Polynomial Time: From Randomized Algorithms to “PRIMES is in P”, volume 3000 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-40344-2. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 147 pp. LCCN QA246. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t3000.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40344-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3000;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12334.

Dignum:2004:AAC

[Dig04] Frank Dignum, editor. Advances in Agent Communication: International Workshop on Agent Communication Languages, ACL 2003, Melbourne, Australia, July 14, 2003: Revised and Invited Papers, volume 2922 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20769-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 I5925 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2922.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20769-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2922;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=

volume&id=doi:10.1007/
b94813.

Dignum:2011:AGS

[Dig11] Frank Dignum, editor.

Agents for Games and Simulations II: Trends in Techniques, Concepts and Design, volume 6525 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18180-5 (print), 3-642-18181-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18181-8.

Dikaiakos:2004:GCS

[Dik04] Marios D. Dikaiakos, editor. Grid computing: Second European AcrossGrids Conference, AxGrids 2004, Nicosia, Cyprus, January 28– 30, 2004: revised papers, volume 3165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22888-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 E97 2004. URL http:

//www.springerlink.
com/content/978-3-540-

                                        22888-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3165;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99982.

Dill:1994:CAV

[Dil94] David L. Dill, editor. Computer aided verification: 6th international conference, CAV ’94, Stanford, California, USA, June 21–23, 1994: proceedings, volume 818 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58179-0, 0-387-58179-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 C38 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0818.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58179-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=818.

Dimov:2003:NMA

[Dim03] Ivan Dimov, editor. Numerical methods and applications: 5th International Conference, NMA 2002, Borovets, Bulgaria, August 20– 24, 2002: Revised Papers, volume 2542 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00608-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA297 .I483 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2542.htm;
 http://www.springerlink.
com/content/978-3-540-
00608-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2542.

Also available via the World Wide Web.

Dediu:2009:LAT

[DIMV09] Adrian Horia Dediu, Armand Mihai Ionescu, and Carlos Martín-Vide, editors. Language and Automata Theory and Applications: Third International Conference, LATA 2009, Tarragona, Spain, April 2–8, 2009. Proceedings, volume 5457 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00981-6 (print), 3-642-00982-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00982-2.

Dediu:2011:LAT

[DIMV11] Adrian-Horia Dediu, Shunsuke Inenaga, and Carlos Martin-Vide, editors. Language and automata theory and applications: 5th international conference, LATA 2011, Tarragona, Spain, May 26–31, 2011. Proceedings, volume 6638 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2011. ISBN 3-642-21253-0 (softcover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21253-6.

Dingledine:2003:PETb

[Din03] Roger Dingledine, editor. Privacy Enhancing Technologies: Third International Workshop, PET 2003, Dresden, Germany, March 26– 28, 2003: Revised Papers, volume 2760 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20610-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 P49 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2760.
htm;   http://www.springerlink.
com/content/978-3-540-
20610-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2760;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94512.

Dingsoyr:2004:SPI

[Din04] Torgeir Dingsøyr, editor. Software process improvement: 11th European conference, EuroSPI 2004, Trondheim, Norway, November 10– 12, 2004: proceedings, volume 3281 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23725-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 E77 2004.

URL http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3281;
http://www.springerlink.
com/content/978-3-540-
23725-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3281;
http://www.springerlink.
com/openurl.asp?genre=

volume&id=doi:10.1007/
b102170.

Domenach:2012:FCA

[DIP12] Florent Domenach, Dmitry I. Ignatov, and Jonas Poelmans, editors. Formal Concept Analysis: 10th International Conference, ICFCA 2012, Leuven, Belgium, May 7–10, 2012. Proceedings, volume 7278 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29891-5 (print), 3-642-29892-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29892-9.

Dehne:2011:ADS

[DIS11] Frank Dehne, John Iacono, and Jörg-Rüdiger Sack, editors. Algorithms and Data Structures: 12th International Symposium, WADS 2011, New York, NY, USA, August 15–17, 2011. Proceedings, volume 6844 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22299-4 (print), 3-642-22300-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22300-6.

Dittrich:1988:AOO

[Dit88] Klaus R. Dittrich, editor. Advances in objectoriented database systems: 2nd International Workshop on Objectoriented Database Systems, Bad Münster am Stein-Ebernburg, FRG, September 27–30, 1988, proceedings, volume 334 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-50345-5 (USA) (Germany, West). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I68 1988. DM50.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0334.htm;
http://www.springerlink.
com/content/978-0-387-
50345-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=334.

Dittrich:2001:ODI

[Dit01] Klaus R. Dittrich, editor. Objects and databases: International Symposium, Sophia Antipolis, France, June 13, 2000: Revised Papers, volume 1944 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41664-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .O285 2001; QA267.A1 L43 no.1944. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1944.htm;
 http://www.springerlink.
com/content/978-3-540-
41664-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1944.

Deransart:1990:AGT

[DJ90] Pierre Deransart and Martin Jourdan, editors. Attribute grammars and their applications: international conference WAGA, Paris, France, September 19–21, 1990: proceedings, volume 461 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-53101-7 (Berlin), 0-387-53101-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .A78 1990. DM52.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0461.
htm;   http://www.springerlink.
com/content/978-3-540-
53101-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=461.

Proceedings of the International Workshop on Attribute Grammars and Their Applications (WAGA) organized by the Institut national de recherche en informatique et en automatique, sponsored by the Association française pour la cybernetique économique et technique.

Doermann:2008:ACH

[DJ08] David Doermann and Stefan Jaeger, editors. Arabic and Chinese Handwriting Recognition: SACH 2006 Summit College Park, MD, USA, September 27–28, 2006 Selected Papers, volume 4768 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78198-6 (print), 3-540-78199-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        78199-8.

Davies:2007:ICA

[DJAP07] Mike E. Davies, Christopher J. James, Samer A. Abdallah, and Mark D. Plumbley, editors. Independent Component Analysis and Signal Separation: 7th International Conference, ICA 2007, London, UK, September 9–12, 2007. Proceedings, volume 4666 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74493-2 (print), 3-540-74494-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74494-8.

Djidjev:1989:OAI

[Dji89] H. Djidjev, editor. Optimal algorithms: International Symposium, Varna, Bulgaria, May 29–June 2 1989 proceedings, volume 401 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51859-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA297.5 .O66 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0401.htm;
 http://www.springerlink.
com/content/978-0-387-
51859-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=401.

Papers presented at the Second International Symposium on Optimal Algorithms held May 29– June 2, 1989, in Varna, Bulgaria, sponsored by the Center of Informatics and Computer Technology of the Bulgarian Academy of Sciences — Frwd.

Dieste:2012:PFS

[DJJ12] Oscar Dieste, Andreas Jedlitschka, and Natalia Juristo, editors. Product-Focused Software Process Improvement: 13th International Conference, PROFES 2012, Madrid, Spain, June 13–15, 2012 Proceedings, volume 7343 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31062-1 (print), 3-642-31063-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        31063-8.

Deransart:1988:AGD

[DJL88] Pierre Deransart, Martin Jourdan, and Bernard Lorho. Attribute grammars: definitions, systems, and bibliography, volume 323 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-50056-1 (USA), 3-540-50056-1. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 232 pp. LCCN QA76.7 .D461 1988. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0323.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        50056-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=323.

Detyniecki:2006:AMR

[DJNR06] Marcin Detyniecki, Joemon M. Jose, Andreas Nürnberger, and C. J. ‘Keith’ Rijsbergen. Adaptive Multimedia Retrieval: User, Context, and Feedback: Third International Workshop, AMR 2005, Glasgow, UK, July 28–29, 2005, Revised Selected Papers, volume 3877 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32174-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32174-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3877.

Dinur:2009:ARC

[DJNR09] Irit Dinur, Klaus Jansen, Joseph Naor, and José Rolim, editors. Approximation, randomization, and combinatorial optimization: algorithms and techniques: 12th International Workshop, AP-PROX 2009, and 13th International Workshop, RANDOM 2009, Berkeley, CA, USA, August, 21–23, 2009: proceedings, volume 5687 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-642-03684-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I643 2009. URL

http://www.loc.gov/
catdir/enhancements/
fy1208/2009932147-d.
html;   http://www.

                                        springerlink.com/content/
                                        978-3-642-03684-2;
                                        http://www.springerlink.
                                        com/content/978-3-642-
                                        03684-2/.

Diaz:2006:ARC

[DJRZ06] J. (Josep) Díaz, Klaus Jansen, José D. P. Rolim, and Uri Zwick, editors. Approximation, randomization and combinatorial optimization: algorithms and techniques: 9th International Workshop on Approximation Algorithms for Combinatorial Optimization Problems, APPROX 2006, and 10th International Workshop on Randomization and Computation, RANDOM 2006, Barcelona, Spain, August 28–30, 2006: proceedings, volume 4110 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-38044-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I643 2006. URL

                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0825/2006931401-d.
                                        html;   http://www.
                                        loc.gov/catdir/toc/
                                        fy0703/2006931401.html;
                                         http://www.springerlink.
                                        com/content/978-3-540-

38044-3;   http://www.
springerlink.com/content/
978-3-540-38044-3/;
http://www.springerlink.
com/openurl.asp?genre=
issue\%26issn=0302-
9743\%26volume=4110.

Dix:1991:NIL

[DJS91] J. Dix, K. P. Jantke, and P. H. Schmitt, editors. Nonmonotonic and inductive logic: 1st international workshop, Karlsruhe, Germany, December 4–7, 1990, proceedings, volume 543 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54564-6, 0-387-54564-

6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .N66 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0543.htm;
http://www.springerlink.
com/content/978-3-540-
54564-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=543.

Proceedings of the First International Workshop on Nonmonotonic and Inductive Inference.

Despotovic:2006:APP

[DJS06] Zoran Despotovic, Sam Joseph, and Claudio Sartori, editors. Agents and Peer-to-Peer Computing: 4th International Workshop, AP2PC 2005, Utrecht, The Netherlands, July 25, 2005. Revised Papers, volume 4118 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49025-6 (print), 3-540-68967-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68967-6.

Dengel:2004:RLA

[DJW04] Andreas Dengel, Markus Junker, and Anette Weisbecker, editors. Reading and Learning: Adaptive Content Recognition, volume 2956 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21904-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T58.64 .R43 2004. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t2956.htm;
http://www.springerlink.
com/content/978-3-540-
21904-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2956;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96988.

Dassow:1989:MLC

[DK89] Jurgen Dassow and Jozef Kelemen, editors. Machines, languages, and complexity, volume 381 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51516-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I57731 1988. US$21.40. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0381.htm;
http://www.springerlink.
com/content/978-0-387-
51516-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=381.

Dassow:1990:APT

[DK90] Jurgen Dassow and Jozef Kelemen, editors. Aspects and prospects of theoretical computer science: 6th International Meeting of Young Computer Scientists, Smolenice, Czechoslovakia, November 19–23, 1990: proceedings, volume 464 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-53414-8 (USA), 3-540-53414-8 (Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I628 1990. DM47.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0464.
htm;   http://www.springerlink.
com/content/978-0-387-
53414-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=464.

Dohi:2002:MICa

[DK02a] Takeyoshi Dohi and Ron Kikinis, editors. Medical Image Computing and Computer-Assisted Intervention— MICCAI 2002: 5th International Conference, Tokyo, Japan, September 25–28, 2002, Proceedings, Part I, volume 2488 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44224-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN WB 141 I605m 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2488.htm;
 http://www.springerlink.
com/content/978-3-540-
44224-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2488.

Dohi:2002:MICb

[DK02b] Takeyoshi Dohi and Ron Kikinis, editors. Medical Image Computing and Computer-Assisted Intervention— MICCAI 2002: 5th International Conference, Tokyo, Japan, September 25–28, 2002, Proceedings, Part II, volume 2489 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44225-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN WB 141 I605m 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2489.htm;
 http://www.springerlink.
com/content/978-3-540-
44225-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2489.

Dojat:2003:AIM

[DKB03] Michel Dojat, Elpida Keravnou, and Pedro Barahona, editors. Artificial Intelligence in Medicine: 9th Conference on Artificial Intelligence in Medicine in Europe, AIME 2003, Protaras, Cyprus, October 18–22, 2003 proceedings, volume 2780 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20129-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R859.7.A78 C66 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2780.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20129-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2780;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12009.

DeBra:2010:UMA

[DKC10] Paul De Bra, Alfred Kobsa, and David Chin, editors. User Modeling, Adaptation, and Personalization: 18th International Conference, UMAP 2010, Big Island, HI, USA, June 20–24, 2010. Proceedings, volume 6075 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13469-6 (print), 3-642-13470-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13470-8.

DiMartino:2005:RAP

[DKD05] Beniamino Di Martino, Dieter Kranzlmüller, and J. J. Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 12th European PVM/MPI User’s Group Meeting, Sorrento, Italy, September 18–21, 2005: proceedings, volume 3666 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29009-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3666;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        29009-4.

deKort:2007:PTS

[dKIM^{+}07] Yvonne de Kort, Wijnand IJsselsteijn, Cees Midden, Berry Eggen, and B. J. Fogg, editors. Persuasive Technology: Second International Conference on Persuasive Technology, PERSUASIVE 2007, Palo Alto, CA, USA, April 26–27, 2007, Revised Selected Papers, volume 4744 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77005-4 (print), 3-540-77006-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77006-0.

DeTurck:2008:MLS

[DKK08] Filip De Turck, Wolfgang Kellerer, and George Kormentzas, editors. Managing Large-Scale Service Deployment: 19th IFIP/ IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2008, Samos Island, Greece, September 22– 26, 2008. Proceedings, volume 5273 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85999-3 (print), 3-540-87353-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87353-2.

Diaz:2004:ALP

[DKLS04] Josep Díaz, Juhani Karhumäki, Arto Lepistö, and Donald Sannella, editors. Automata, Languages and Programming: 31st International Colloquium, ICALP 2004, Turku, Finland, July 12–16, 2004: Proceedings, volume 3142 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22849-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3142.htm;
 http://www.springerlink.
com/content/978-3-540-
22849-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=3142;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99859.

Dey:2005:MUC

[DKLT05] Anind Dey, Boicho Kokinov, David Leake, and Roy Turner, editors. Modeling and Using Context: 5th International and Interdisciplinary Conference CON-TEXT 2005, Paris, France, July 5–8, 2005. Proceedings, volume 3554 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26924-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65 I537 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        26924-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3554.

Delobel:1991:DOO

[DKM91] C. Delobel, M. Kifer, and Y. Masunaga, editors. Deductive and objectoriented databases: Second International Conference, DOOD ’91, Munich, Germany, December 16–18, 1991: proceedings, volume 566 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-55015-1, 0-387-55015-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55826 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0566.
htm;   http://www.springerlink.
com/content/978-3-540-
55015-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=566.

Delcambre:2005:CMI

[DKM^{+}05] Lois Delcambre, Christian Kop, Heinrich C. Mayr, John Mylopoulos, and Oscar Pastor, editors. Conceptual Modeling — ER 2005: 24th International Conference on Conceptual Modeling, Klagenfurt, Austria, October 24–28, 2005. Proceedings, volume 3716 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29389-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29389-7;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3716.

Dunin-Keplicz:2002:TPM

[DKN02] Barbara Dunin-Keplicz and Edward Nawarecki, editors. From theory to practice in multi-agent systems: Second International Workshop of Central and Eastern Europe on Multi-Agent Systems, CEEMAS 2001, Cracow, Poland, September 26– 29, 2001: Revised Papers, volume 2296 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43370-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 C43 2001.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2296.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        43370-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2296.

Detyniecki:2011:AMRb

[DKN^{+}11] Marcin Detyniecki, Peter Knees, Andreas Nürnberger, Markus Schedl, and Sebastian Stober, editors. Adaptive Multimedia Retrieval. Context, Exploration, and Fusion: 8th International Workshop, AMR 2010, Linz, Austria, August 17–18, 2010, Revised Selected Papers, volume 6817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-27168-5 (print), 3-642-27169-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27169-4.

Dinneen:2012:CPB

[DKN12] Michael J. Dinneen, Bakhadyr Khoussainov, and André Nies, editors. Computation, Physics and Beyond: International Workshop on Theoretical Computer Science, WTCS 2012, Dedicated to Cristian S. Calude on the Occasion of His 60th Birthday, Auckland, New Zealand, February 21– 24, 2012, Revised Selected and Invited Papers, volume 7160 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27653-9 (print), 3-642-27654-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27654-5.

Duch:2005:ANNb

[DKOZ05a] Włodzisław Duch, Janusz Kacprzyk, Erkki Oja, and Sławomir Zadrozny, editors. Artificial neural networks: Biological Inspirations — ICANN 2005: 15th International Conference, Warsaw, Poland, September 11–15, 2005: proceedings, volume 3696 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28752-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28752-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3696.

Duch:2005:ANNa

[DKOZ05b] Włodzisław Duch, Janusz Kacprzyk, Erkki Oja, and Sławomir Zadrozny, editors. Artificial neural networks: Formal Models and Their Applications

— ICANN 2005: 15th International Conference, Warsaw, Poland, September 11–15, 2005: proceedings, volume 3697 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28755-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28755-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3697.

Dongarra:2000:RAP

[DKP00] J. J. Dongarra, Péter Kacsuk, and Norbert Podhorszki, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 7th European PVM/MPI Users’ Group Meeting, Balatonfüred, Hungary, September 10– 13, 2000. Proceedings, volume 1908 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41010-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973 2000. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t1908.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41010-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1908.

DalCin:2005:DCE

[DKP05] Mario Dal Cin, Mohamed Kaâniche, and András Pataricza, editors. Dependable computing: EDCC 2005: 5th European dependable computing conference, Budapest, Hungary, April 20–22, 2005, proceedings, volume 3463 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25723-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25723-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3463;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b107276.

Druschel:2002:PPS

[DKR02] Peter Druschel, Frans Kaashoek, and Antony Rowstron, editors. Peerto-Peer Systems: First International Workshop, IPTPS 2002 Cambridge, MA, USA, March 7– 8, 2002. Revised Papers, volume 2429 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44179-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.525 .I68 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2429.
htm;   http://www.springerlink.
com/content/978-3-540-
44179-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2429.

Dony:2006:ATE

[DKRT06] Christophe Dony, Jørgen Lindskov Knudsen, Alexander Romanovsky, and Anand Tripathi, editors. Advanced Topics in Exception Handling Techniques, volume 4119 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37443-4 (print), 3-540-37445-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        37445-9.

DeBosschere:2007:HPE

[DKS^{+}07] Koen De Bosschere, David Kaeli, Per Stenström, David Whalley, and Theo Ungerer, editors. High Performance Embedded Architectures and Compilers: Second International Conference, HiPEAC 2007, Ghent, Belgium, January 28–30, 2007. Proceedings, volume 4367 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69337-8 (print), 3-540-69338-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69338-3.

Dusterhoft:2012:CMT

[DKS12] Antje Düsterhöft, Meike Klettke, and Klaus-Dieter Schewe, editors. Conceptual Modelling and Its Theoretical Foundations: Essays Dedicated to Bernhard Thalheim on the Occasion of His 60th Birthday, volume 7260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28278-4 (print), 3-642-28279-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28279-9.

Dittmann:2005:CMS

[DKU05] Jana Dittmann, Stefan Katzenbeisser, and Andreas Uhl, editors. Communications and multimedia security: 9th IFIP TC-6 TC-11 International Conference, CMS 2005, Salzburg, Austria, September 19–21, 2005: proceedings, volume 3677 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28791-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28791-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3677.

Derniame:1999:SPP

[DKW99] Jean-Claude Derniame, Badara Ali Kaba, and David Wastell, editors. Software Process: Principles, Methodology, and Technology, volume 1500 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65516-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 307 pp. LCCN QA76.76.D47 S6638 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1500.
htm;   http://www.springerlink.
com/content/978-3-540-
65516-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1500.

Duval:2007:CNL

[DKW07] Erik Duval, Ralf Klamma, and Martin Wolpers, editors. Creating New Learning Experiences on a Global Scale: Second European Conference on Technology Enhanced Learning, EC-TEL 2007, Crete, Greece, September 17–20, 2007. Proceedings, volume 4753 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75194-7 (print), 3-540-75195-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
75195-3.

Dalkilic:2006:DMB

[DKY06] Mehmet M. Dalkilic, Sun Kim, and Jiong Yang, editors. Data Mining and Bioinformatics: First International Workshop, VDMB 2006, Seoul, Korea, September 11, 2006, Revised Selected Papers, volume 4316 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68970-2 (print), 3-540-68971-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68971-3.

Duijvestijn:1981:TIP

[DL81] A. J. W. Duijvestijn and P. C. Lockemann, editors. Trends in information processing systems: 3rd conference of the European Cooperation in Informatics, Munich, October 20–22, 1981: proceedings, volume 123 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10885-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .T73. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0123.
htm;   http://www.springerlink.
com/content/978-0-387-
10885-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=123.

Held in conjunction with the 11th annual conference of the Gesellschaft für Informatik.

Dershowitz:1995:CTR

[DL95a] Nachum Dershowitz and Naomi Lindenstrauss, editors. Conditional and typed rewriting systems: 4th International Workshop, CTRS-94, Jerusalem, Israel, July 13–15, 1994: proceedings, volume 968 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60381-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I5783 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0968.htm;
 http://www.springerlink.
com/content/978-3-540-
60381-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

0302-9743&volume=968.

Held in conjunction with ICALP 94.

Du:1995:CCF

[DL95b] Dingzhu Du and Ming Li, editors. Computing and combinatorics: first annual international conference, CO-COON’95, Xi’an, China, August 1995: proceedings, volume 959 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60216-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I5443 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0959.htm;
http://www.springerlink.
com/content/978-3-540-
60216-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=959.

Demoen:2004:LPI

[DL04a] Bart Demoen and Vladimir Lifschitz, editors. Logic programming: 20th international conference, ICLP 2004, Saint-Malo, France, September 6– 10, 2004, proceedings, volume 3132 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22671-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
22671-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3132;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99475.

Dix:2004:CLM

[DL04b] J. (Jürgen) Dix and João Alexandre Leite, editors. Computational logic in multi-agent systems: 4th International Workshop, CLIMA IV, Fort Lauderdale, FL., USA, January 6–7, 2004: revised selected and invited papers, volume 3259 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24010-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.L63 C65 2004g.

URL http://www.springerlink.
com/content/978-3-540-
24010-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

0302-9743&volume=3259;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104175.

Damiani:2006:DAS

[DL06a] Ernesto Damiani and Peng Liu, editors. Data and Applications Security XX: 20th Annual IFIP WG 11.3 Working Conference on Data and Applications Security, Sophia Antipolis, France, July 31–August 2, 2006. Proceedings, volume 4127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36796-9 (print), 3-540-36799-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36799-4.

Dan:2006:SOC

[DL06b] Asit Dan and Winfried Lamersdorf, editors. Service-Oriented Computing — ICSOC 2006: 4th International Conference, Chicago, IL, USA, December 4–7, 2006. Proceedings, volume 4294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68147-7 (print), 3-540-68148-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68148-9.

Dwyer:2007:FAS

[DL07] Matthew B. Dwyer and Antónia Lopes, editors. Fundamental Approaches to Software Engineering: 10th International Conference, FASE 2007, Held as Part of the Joint European Conferences, on Theory and Practice of Software, ETAPS 2007, Braga, Portugal, March 24 — April 1, 2007. Proceedings, volume 4422 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71288-7 (print), 3-540-71289-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71289-3.

delaBanda:2008:LPI

[dlBP08] Maria Garcia de la Banda and Enrico Pontelli, editors. Logic Programming: 24th International Conference, ICLP 2008 Udine, Italy, December 9–13 2008 Proceedings, volume 5366 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89981-2 (print), 3-540-89982-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89982-2.

deLemos:2008:ADS

[dLDG^{+}08] Rogério de Lemos, Felicita Di Giandomenico, Cristina Gacek, Henry Muccini, and Marlon Vieira, editors. Architecting Dependable Systems V, volume 5135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85570-X (print), 3-540-85571-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85571-2.

Dini:2004:SAP

[DLdS04] Petre Dini, Pascal Lorenz, and José Neuman de Souza, editors. Service Assurance with Partial and Intermittent Resources: First International Workshop, SAPIR 2004, Fortaleza, Brazil, August 1–6, 2004: Proceedings, volume 3126 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22567-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.75 .S264 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3126.
htm;   http://www.springerlink.
com/content/978-3-540-
22567-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3126;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99076.

deLemos:2009:ADS

[dLFG^{+}09] Rogério de Lemos, Jean-Charles Fabre, Cristina Gacek, Fabio Gadducci, and Maurice ter Beek, editors. Architecting Dependable Systems VI, volume 5835 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10247-6 (print), 3-642-10248-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10248-6.

DInverno:2002:FAM

[DLFP02] Mark D’Inverno, Michael Luck, Michael Fisher, and Chris Preist, editors. Foundations and Applications of Multi-Agent Systems: UKMAS Workshop 1996-2000. Selected Papers, volume 2403 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43962-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 U36 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2403.htm;
http://www.springerlink.
com/content/978-3-540-
43962-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2403.

Dix:2010:CLMb

[DLGJ10] Jürgen Dix, João Leite, Guido Governatori, and Wojtek Jamroga, editors. Computational Logic in Multi-Agent Systems: 11th International Workshop, CLIMA XI, Lisbon, Portugal, August 16–17, 2010. Proceedings, volume 6245 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14976-6 (print), 3-642-14977-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14977-1.

deLemos:2003:ADS

[dLGR03] Rogério de Lemos, Cristina Gacek, and Alexander Romanovsky, editors. Architecting dependable systems, volume 2677 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40727-8. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 307 pp. LCCN QA76.9.A73 A715 2003; QA267.A1 L43. URL http:

//www.springerlink.
com/content/978-3-540-
40727-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2677.

deLemos:2004:ADS

[dLGR04] Rogério de Lemos, Cristina Gacek, and Alexander Romanovsky, editors. Architecting dependable systems II, volume 3069 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23168-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 350 pp. LCCN QA76.8.A73 A73 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23168-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3069;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100023.

deLemos:2005:ADS

[dLGR05] Rogério de Lemos, Cristina Gacek, and Alexander Romanovsky, editors. Architecting Dependable Systems III, volume 3549 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28968-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28968-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3549.

deLemos:2007:ADS

[dLGR07] Rogério de Lemos, Cristina Gacek, and Alexander Romanovsky, editors. Architecting Dependable Systems IV, volume 4615 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74033-3 (print), 3-540-74035-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74035-3.

Durand-Lose:2012:UCN

[DLJ12] Jérôme Durand-Lose and Nataša Jonoska, editors. Unconventional Computation and Natural Computation: 11th International Conference, UCNC 2012, Orléan, France, September 3– 7, 2012. Proceedings, volume 7445 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32893-8 (print), 3-642-32894-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32894-7.

Deransart:1989:PLI

[DLM89] Pierre Deransart, Bernard Lorho, and J. Maluszynski. Programming languages implementation and logic programming, volume 348 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-50820-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 299 pp. LCCN QA76.7 .I591 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0348.htm;
http://www.springerlink.
com/content/978-0-387-
50820-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=348.

Dongarra:1999:RAP

[DLM99] J. J. Dongarra, E. Luque, and Tomas Margalef, editors. Recent advances in parallel virtual machine and message passing interface: 6th European PVM/ MPI Users’ Group Meeting, Barcelona, Spain, September 26–29, 1999: Proceedings, volume 1697 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1697.htm;
 http://www.springerlink.
com/content/978-3-540-
66549-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1697.

Durand-Lose:2007:MCU

[DLM07] Jérôme Durand-Lose and Maurice Margenstern, editors. Machines, Computations, and Universality: 5th International Conference, MCU 2007, Orléans, France, September 10–13, 2007. Proceedings, volume 4664 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74592-0 (print), 3-540-74593-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
74593-8.

Detyniecki:2010:AMR

[DLN10] Marcin Detyniecki, Ulrich Leiner, and Andreas Nürnberger, editors. Adaptive Multimedia Retrieval. Identifying, Summarizing, and Recommending Image and Music: 6th International Workshop, AMR 2008, Berlin, Germany, June 26–27, 2008. Revised Selected Papers, volume 5811 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14757-7 (print), 3-642-14758-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14758-6.

DeDecker:2011:CMS

[DLNU11] Bart De Decker, Jorn Lapon, Vincent Naessens, and Andreas Uhl, editors. Communications and Multimedia Security: 12th IFIP TC 6 /TC 11 International Conference, CMS 2011, Ghent, Belgium, October 19–21,2011. Proceedings, volume 7025 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24711-3 (print), 3-642-24712-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24712-5.

Dongarra:2003:RAP

[DLO03] Jack Dongarra, Domenico Laforenza, and Salvatore Orlando, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 10th European PVM/MPI User’s group Meeting, Venice, Italy, September 29–October 2, 2003: Proceedings, volume 2840 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20149-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2840.htm;
 http://www.springerlink.
com/content/978-3-540-
20149-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2840.

Dehne:2005:ADS

[DLOS05] Frank Dehne, Alejandro López-Ortiz, and Jörg-Rüdiger Sack, editors. Algorithms and data structures: 9th international workshop, WADS 2005, Waterloo, Canada, August 15–17, 2005: proceedings, volume 3608 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28101-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 W33 2005.

URL http://www.springerlink.
com/content/978-3-540-
28101-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3608.

Dobnikar:2011:ANCa

[DLŠ11a] Andrej Dobnikar, Uroš Lotrič, and Branko Šter, editors. Adaptive and Natural Computing Algorithms: 10th International Conference, ICANNGA 2011, Ljubljana, Slovenia, April 14– 16, 2011, Proceedings, Part I, volume 6593 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20281-0 (print), 3-642-20282-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20282-7.

Dobnikar:2011:ANCb

[DLŠ11b] Andrej Dobnikar, Uroš Lotrič, and Branko Šter, editors. Adaptive and Natural Computing Algorithms: 10th International Conference, ICANNGA 2011, Ljubljana, Slovenia, April 14– 16, 2011, Proceedings, Part II, volume 6594 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20266-7 (print), 3-642-20267-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20267-4.

Davidsson:2005:MAM

[DLT05] Paul Davidsson, Brian Logan, and Keiki Takadama, editors. Multi-agent and multi-agent-based simulation: joint workshop MABS 2004, New York, NY, USA, July 19, 2004: Revised selected papers, volume 3415 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25262-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M32 2005. URL http:

//www.springerlink.
com/content/978-3-540-
25262-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3415;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106991.

Danelutto:2004:EPP

[DLV04] Marco Danelutto, Domenico Laforenza, and Marco Vanneschi, editors. Euro-Par 2004 parallel processing: 10th International Euro-Par Conference, Pisa, Italy, August 31–September 3, 2004: proceedings, volume 3149 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22924-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I554 2004. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t3149.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22924-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3149.

Dosch:2006:SER

[DLW06] Walter Dosch, Roger Y. Lee, and Chisu Wu, editors. Software engineering research and applications: second international conference, SERA 2004, Los Angeles, CA, USA, May 5– 7, 2004, revised selected papers, volume 3647 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32133-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32133-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3647.

Dong:2007:ADW

[DLW^{+}07] Guozhu Dong, Xuemin Lin, Wei Wang, Yun Yang, and Jeffrey Xu Yu, editors. Advances in Data and Web Management: Joint 9th Asia-Pacific Web Conference, APWeb 2007, and 8th International Conference, on Web-Age Information Management, WAIM 2007, Huang Shan, China, June 16– 18, 2007. Proceedings, volume 4505 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72483-4 (print), 3-540-72524-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72524-4.

Desai:2012:PPM

[DLW12] Nirmit Desai, Alan Liu, and Michael Winikoff, editors. Principles and Practice of Multi-Agent Systems: 13th International Conference, PRIMA 2010, Kolkata, India, November 12–15, 2010, Revised Selected Papers, volume 7057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25919-7 (print), 3-642-25920-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        25920-3.

deLemos:2003:DCL

[dLWC03] Rogerio de Lemos, Taisy Silva Weber, and Joäo Batista Camargo, Jr., editors. Dependable Computing Latin-American symposium, LADC 2003, São Paulo, Brazil, October 21–24, 2003: Proceedings, volume 2847 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20224-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F3. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2847.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20224-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2847;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13952.

deLara:2012:FAS

[dLZ12] Juan de Lara and Andrea Zisman, editors. Fundamental Approaches to Software Engineering: 15th International Conference, FASE 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7212 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28871-5 (print), 3-642-28872-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28872-2.

Deransart:1990:PLI

[DM90] Pierre Deransart and J. Małuszynski. Programming language implementation and logic programming: proceedings, volume 456 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 3-540-53010-X, 0-387-53010-X. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 401 pp. LCCN QA76.7 .I59 1990. DM57.50. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0456.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        53010-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=456.

Davidson:2001:LCT

[DM01a] Jack W. Davidson and Sang Lyul Min, editors. Languages, compilers, and tools for embedded systems: ACM SIGPLAN Workshop LCTES 2000, Vancouver, Canada, June 18, 2000: Proceedings, volume 1985 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41781-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .A13 2000; QA267.A1 L43 no.1985. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1985.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41781-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1985.

Davidsson:2001:MAB

[DM01b] Paul Davidsson and Scott J. Moss, editors. Multi-agent-based simulation: Second International Workshop, MABS 2000, Boston, MA, USA, July: Revised Papers, volume 1979 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41522-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M32 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1979.
htm;   http://www.springerlink.
com/content/978-3-540-
41522-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1979.

Danezis:2006:PETa

[DM06] George Danezis and David Martin, editors. Privacy Enhancing Technologies: 5th International Workshop, PET 2005, Cavtat, Croatia, May 30–June 1, 2005, Revised Selected Papers, volume 3856 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34745-3 (print), 3-540-34746-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        34746-0.

deMeer:2005:QSI

[dMB05] Hermann de Meer and Nina Bhatti, editors. Quality of Service — IWQoS 2005: 13th International Workshop, IWQoS 2005, Passau, Germany, June 21–23, 2005. Proceedings, volume 3552 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26294-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5956 .I57 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        26294-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3552.

Diniz:2007:RCA

[DMB^{+}07] Pedro C. Diniz, Eduardo Marques, Koen Bertels, Marcio Merino Fernandes, and João M. P. Cardoso, editors. Reconfigurable Computing: Architectures, Tools and Applications: Third International Workshop, ARC 2007, Mangaratiba, Brazil, March 27– 29, 2007. Proceedings, volume 4419 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71430-8 (print), 3-540-71431-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71431-6.

DeTroyer:2011:ACM

[DMB^{+}11] Olga De Troyer, Claudia Bauzer Medeiros, Roland Billen, Pierre Hallot, Alkis Simitsis, and Hans Van Mingroot, editors. Advances in Conceptual Modeling. Recent Developments and New Directions: ER 2011 Workshops FP-UML, MoRE-BI, Onto-CoM, SeCoGIS, Variability@ER, WISM, Brussels, Belgium, October 31 — November 3, 2011. Proceedings, volume 6999 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24573-0 (print), 3-642-24574-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24574-9.

deMoor:2011:DRF

[dMGFS11] Oege de Moor, Georg Gottlob, Tim Furche, and Andrew Sellers, editors. Datalog Reloaded: First International Workshop, Datalog 2010, Oxford, UK, March 16–19, 2010. Revised Selected Papers, volume 6702 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24205-7 (print), 3-642-24206-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24206-9.

deMoor:2003:CSK

[dMLG03] Aldo de Moor, Wilfried Lex, and Bernhard Ganter, editors. Conceptual Structures for Knowledge Creation and Communication: 11th International Conference on Conceptual Structures, ICCS 2003, Dresden, Germany, July 21– 25, 2003: Proceedings, volume 2746 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40576-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2 .I58 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2746.htm;
http://www.springerlink.
com/content/978-3-540-
40576-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2746;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11835.

Deng:2011:AICa

[DMLW11a] Hepu Deng, Duoqian Miao, Jingsheng Lei, and Fu Lee Wang, editors. Artificial Intelligence and Computational Intelligence: Third International Conference, AICI 2011, Taiyuan, China, September 24– 25, 2011, Proceedings, Part I, volume 7002 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23880-7 (print), 3-642-23881-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23881-9.

Deng:2011:AICb

[DMLW11b] Hepu Deng, Duoqian Miao, Jingsheng Lei, and Fu Lee Wang, editors. Artificial Intelligence and Computational Intelligence: Third International Conference, AICI 2011, Taiyuan, China, September 24– 25, 2011, Proceedings, Part II, volume 7003 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23886-6 (print), 3-642-23887-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23887-1.

Deng:2011:AICc

[DMLW11c] Hepu Deng, Duoqian Miao, Jingsheng Lei, and Fu Lee Wang, editors. Artificial Intelligence and Computational Intelligence: Third International Conference, AICI 2011, Taiyuan, China, September 24– 25, 2011, Proceedings, Part III, volume 7004 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23895-5 (print), 3-642-23896-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23896-3.

Davida:1998:ISF

[DMO98] George Davida, Masahiro Mambo, and Eiji Okamoto, editors. Information security: First International Workshop, ISW ’97, Tatsunokuchi, Ishikawa, Japan, September 17– 19, 1997: Proceedings, volume 1396 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64382-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25I89 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1396.htm;
http://www.springerlink.
com/content/978-3-540-
64382-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1396.

Daniilidis:2010:CVEa

[DMP10a] Kostas Daniilidis, Petros Maragos, and Nikos Paragios, editors. Computer Vision — ECCV 2010: 11th European Conference on Computer Vision, Heraklion, Crete, Greece, September 5–11, 2010, Proceedings, Part I, volume 6311 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15548-0 (print), 3-642-15549-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15549-9.

Daniilidis:2010:CVEb

[DMP10b] Kostas Daniilidis, Petros Maragos, and Nikos Paragios, editors. Computer Vision — ECCV 2010: 11th European Conference on Computer Vision, Heraklion, Crete, Greece, September 5–11, 2010, Proceedings, Part II, volume 6312 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15551-0 (print), 3-642-15552-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15552-9.

Daniilidis:2010:CVEc

[DMP10c] Kostas Daniilidis, Petros Maragos, and Nikos Paragios, editors. Computer Vision — ECCV 2010: 11th European Conference on Computer Vision, Heraklion, Crete, Greece, September 5–11, 2010, Proceedings, Part III, volume 6313 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15557-X (print), 3-642-15558-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15558-1.

Daniilidis:2010:CVEd

[DMP10d] Kostas Daniilidis, Petros Maragos, and Nikos Paragios, editors. Computer Vision — ECCV 2010: 11th European Conference on Computer Vision, Heraklion, Crete, Greece, September 5–11, 2010, Proceedings, Part IV, volume 6314 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15560-X (print), 3-642-15561-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15561-1.

Daniilidis:2010:CVEe

[DMP10e] Kostas Daniilidis, Petros Maragos, and Nikos Paragios, editors. Computer Vision — ECCV 2010: 11th European Conference on Computer Vision, Heraklion, Crete, Greece, September 5–11, 2010, Proceedings, Part V, volume 6315 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15554-5 (print), 3-642-15555-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15555-0.

Daniilidis:2010:CVEf

[DMP10f] Kostas Daniilidis, Petros Maragos, and Nikos Paragios, editors. Computer Vision — ECCV 2010: 11th European Conference on Computer Vision, Heraklion, Crete, Greece, September 5–11, 2010, Proceedings, Part VI, volume 6316 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15566-9 (print), 3-642-15567-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15567-3.

DeGroote:2001:LAC

[DMR01] Philippe De Groote, Glyn V. Morrill, and Christian Retore, editors. Logical aspects of computational linguistics: 4th International Conference, LACL 2001, Le Croisic, France, June 27–29, 2001: Proceedings, volume 2099 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42273-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2099; P98.5.L63 L33 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2099.
htm;   http://www.springerlink.
com/content/978-3-540-
42273-0;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2099.

Dimitrakos:2006:FAS

[DMRS06] Theo Dimitrakos, Fabio Martinelli, Peter Y. A. Ryan, and Steve Schneider, editors. Formal Aspects in Security and Trust: Third International Workshop, FAST 2005, Newcastle upon Tyne, UK, July 18–19, 2005, Revised Selected Papers, volume 3866 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32628-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32628-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3866.

Dimitrakos:2007:FAS

[DMRS07] Theo Dimitrakos, Fabio Martinelli, Peter Y. A. Ryan, and Steve Schneider, editors. Formal Aspects in Security and Trust: Fourth International Workshop, FAST 2006, Hamilton, Ontario, Canada, August 26–27, 2006, Revised Selected Papers, volume 4691 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75226-9 (print), 3-540-75227-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75227-1.

Davies:2004:UUC

[DMS04] Nigel Davies, Elizabeth Mynatt, and Itiro Siio, editors. UbiComp 2004: Ubiquitous Computing: 6th International Conference, Nottingham, UK, September 7–10, 2004: Proceedings, volume 3205 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22955-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3205.htm;
 http://www.springerlink.
com/content/978-3-540-
22955-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3205;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b99948.

deMeer:2006:SOS

[dMS06] Hermann de Meer and James P. G. Sterbenz, editors. Self-Organizing Systems: First International Workshop, IW-SOS 2006, and Third International Workshop on New Trends in Network Architectures and Services, EuroNGI 2006, Passau, Germany, September 18–20, 2006 Proceedings, volume 4124 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37658-5 (print), 3-540-37669-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37669-9.

deMoor:2009:CCI

[dMS09] Oege de Moor and Michael I. Schwartzbach, editors. Compiler Construction: 18th International Conference, CC 2009, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2009, York, UK, March 22–29, 2009. Proceedings, volume 5501 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00721-X (print), 3-642-00722-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00722-4.

Drira:2001:CED

[DMV01] Khalil Drira, Andrea Martelli, and Thierry Villemur, editors. Cooperative Environments for Distributed Systems Engineering: The Distributed Systems Environment Report, volume 2236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43083-0. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 279 pp. LCCN QA76.9.D5 C675 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2236.htm;
 http://www.springerlink.
com/content/978-3-540-
43083-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2236.

Dediu:2012:LAT

[DMV12] Adrian-Horia Dediu and Carlos Martín-Vide, editors. Language and Automata Theory and Applications: 6th International Conference, LATA 2012, A Coruña, Spain, March 5–9, 2012. Proceedings, volume 7183 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28331-4 (print), 3-642-28332-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28332-1.

Dongarra:1996:APC

[DMW96] J. J. Dongarra, Kaj Madsen, and Jerzy Waśniewski, editors. Applied parallel computing: computations in physics, chemistry, and engineering science: second International Workshop, PARA ’95, Lyngby, Denmark, August 21–24, 1995: proceedings, volume 1041 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1041.htm;
http://www.springerlink.
com/content/978-3-540-
60902-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1041.

Debar:2000:RAI

[DMW00] Herve Debar, Ludovic Mé, and S. Felix Wu, editors. Recent Advances in Intrusion Detection: Third International Workshop, RAID 2000, Toulouse, France, October 2–4, 2000. Proceedings, volume 1907 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41085-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 R34 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1907.htm;
http://www.springerlink.
com/content/978-3-540-
41085-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1907.

Dongarra:2006:APC

[DMW06] Jack Dongarra, Kaj Madsen, and Jerzy Wasniewski, editors. Applied Parallel Computing: State of the Art in Scientific Computing. 7th International Workshop, PARA 2004, Lyngby, Denmark, June 20–23, 2004. Revised Selected Papers, volume 3732 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-29067-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29067-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3732.

Dierstein:1988:PCS

[DMWW88] Rudiger Dierstein, D. Muller-Wichards, and H.-M. Wacker, editors. Parallel computing in science and engineering: 4th International DFVLR Seminar on Foundations of Engineering Sciences, Bonn, Federal Republic of Germany, June 25–26, 1987, volume 295 of DFVLR seminar series and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-18923-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I549 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0295.htm;
http://www.springerlink.
com/content/978-0-387-
18923-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=295.

Dauchet:1988:CCT

[DN88] M. Dauchet and M. Nivat, editors. CAAP ’88: 13th Colloquium on Trees in Algebra and Programming, Nancy, France, March 21–24, 1988: proceedings, volume 299 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19021-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166.2 .C651 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0299.htm;
http://www.springerlink.
com/content/978-0-387-
19021-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=299.

Held in conjunction with ESOP ’88.

Donatiello:1993:PEC

[DN93] Lorenzo Donatiello and Randolph Nelson, editors. Performance evaluation of computer and communication systems: joint tutorial papers of Performance ’93 and Sigmetrics ’93, volume 729 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57297-X (Berlin), 0-387-57297-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I3428 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0729.htm;
 http://www.springerlink.
com/content/978-3-540-
57297-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=729.

“Tutorial papers presented at the 16th IFIP Working Group 7.3 International Symposium on Computer Performance Modeling, Measurement and Evaluation held in Rome, September 27 to October 1, 1993 and a number of tutorial papers presented at the 1993 ACM Sigmetrics Conference on Measurement and Modeling of Computer Systems held in Santa Clara, California, May 10 to 14, 1993”–Foreword. Contents: Parallel Simulation / R. Ayani — Properties and Analysis of Queueing Network Models with Finite Capacities / S. Balsamo — Performance Analysis and Optimization with the Power-Series Algorithm / H. Blanc — Metropolitan Area Networks (MANs): Protocols, Modeling and Performance Evaluation / M. Conti, E. Gregori and L. Lenzini — Multiprocessor and Distributed System Design: The Integration of Functional Specification and Performance Analysis Using Stochastic Process Algebras / N. Gotz, U. Herzog and M. Rettelbach — Response Time Distributions in Queueing Network Models / P. G. Harrison — Fast Simulation of Rare Events in Queueing and Reliability Models / P. Heidelberger — An Introduction to Modeling Dynamic Behavior with Time Series Analysis / J. L. Hellerstein — Issues in Trace-Driven Simulation / D. R. Kaeli — Maximum Entropy Analysis of Queueing Network Models / D. Kouvatsos — Performance Modeling Using DSPNexpress / C. Lindemann. Contents: Relaxation for Massively Parallel Discrete Event Simulation / B. D. Lubachevsky — The BMAP/G/1 QUEUE: A Tutorial / D. M. Lucantoni — An Overview of Tes Processes and Modeling Methodology / B. Melamed — Performance Engineering of Client-Server Systems / C. Murray Woodside — Queueing Networks with Finite Capacities / R. O. Onvural — Real Time Systems: A Tutorial / F. Panzieri and R. Davoli — Performance Instrumentation Techniques for Parallel Systems / D. A. Reed — A Survey of Bottleneck Analysis in Closed Networks of Queues / P. J. Schweitzer, G. Serazzi and M. Broglia — Software Performance Engineering / C. U. Smith — Performance Measurement Using System Monitors / E. M. Thurner — Providing Quality of Service Packet Switched Networks / D. Towsley — Dependability and Performability Analysis / K. S. Trivedi, G. Ciardo, M. Malhotra and R. A. Sahner — Architectures and Algorithms for Digital Multimedia On-Demand Servers / P. Venkat Rangan. Contents: Analysis and Control of Polling Systems / U. Yechiali — Modeling and Analysis of Transaction Processing Systems / P. S. Yu.

DeBra:2004:AHA

[DN04] Paul De Bra and Wolfgang Nejdl, editors. Adaptive hypermedia and adaptive Web-based systems: Third international conference, AH 2004, Eindhoven, The Netherlands, August 23– 26, 2004, proceedings, volume 3137 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22895-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
22895-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3137;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99480.

Dahl:2007:LPI

[DN07] Véronica Dahl and Ilkka Niemelä, editors. Logic Programming: 23rd International Conference, ICLP 2007, Porto, Portugal, September 8–13, 2007. Proceedings, volume 4670 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74608-0 (print), 3-540-74610-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74610-2.

Diekert:2009:DLT

[DN09] Volker Diekert and Dirk Nowotka, editors. Developments in Language Theory: 13th International Conference, DLT 2009, Stuttgart, Germany, June 30–July 3, 2009. Proceedings, volume 5583 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02736-9 (print), 3-642-02737-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02737-6.

Dybjer:1995:TPP

[DNS95] Peter Dybjer, Bengt Nordstrom, and Jan M. Smith, editors. Types for proofs and programs: International Workshop TYPES ’94, Bastad, Sweden, June 6–10, 1994: proceedings, volume 996 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60579-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I594 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0996.htm;
http://www.springerlink.
com/content/978-3-540-
60579-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=996.

Denzler:2009:PRD

[DNS09] Joachim Denzler, Gunther Notni, and Herbert Süße, editors. Pattern Recognition: 31st DAGM Symposium, Jena, Germany, September 9– 11, 2009. Proceedings, volume 5748 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03797-6 (print), 3-642-03798-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03798-6.

Diaz:1989:TPIa

[DO89a] J. Díaz and F. Orejas, editors. TAP-SOFT ’89: proceedings of the International Joint Conference on Theory and Practice of Software Development, Barcelona, Spain, March 13–17, 1989. Volume 1. Advanced Seminar on Foundations of Innovative Software Development I and Colloquium on Trees in Algebra and Programming (CAAP ’89), volume 351 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 3-540-50939-9 (Berlin: paperback), 0-387-50939-9 (New York: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I57 1989 v. 1–2 (1989). URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t0351.htm;
 http://www.springerlink.
com/content/978-3-540-
50939-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=351.

Diaz:1989:TPIb

[DO89b] J. Díaz and F. Orejas, editors. TAP-SOFT ’89: proceedings of the International Joint Conference on Theory and Practice of Software Development, Barcelona, Spain, March 13–17, 1989. Volume 2, Advanced Seminar on Foundations of Innovative Software Development II and Colloquium on Current Issues in Programming Languages (CCIPL), volume 352 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 3-540-50940-2 (Berlin: paperback), 0-387-50940-2 (New York: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I57 1989 v. 1–2 (1989). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0352.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        50940-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=352.

deOliveira:1998:AAI

[dO98] Flavio Moreira de Oliveira, editor. Advances in artificial intelligence: 14th Brazilian Symposium on Artificial Intelligence, SBIA ’98, Porto Alegre, Brazil, November 4–6, 1998: Proceedings, volume 1515 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65190-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1515. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1515.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65190-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1515.

Damm:2002:FTR

[DO02] Werner Damm and E.-R Olderog, editors. Formal Techniques in Real-Time and Fault-Tolerant Systems: 7th International Symposium, FTRTFT 2002, Co-sponsored by IFIP WG 2.2, Oldenburg, Germany, September 9–12, 2002. Proceedings, volume 2469 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44165-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54 .F75 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2469.htm;
 http://www.springerlink.
com/content/978-3-540-
44165-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2469.

Dolev:2009:OSS

[DO09] Shlomi Dolev and Mihai Oltean, editors. Optical SuperComputing: Second International Workshop, OSC 2009, Bertinoro, Italy, November 18–20, 2009. Proceedings, volume 5882 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10441-X (print), 3-642-10442-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10442-8.

Dolev:2011:OST

[DO11] Shlomi Dolev and Mihai Oltean, editors. Optical Supercomputing: Third International Workshop, OSC 2010, Bertinoro, Italy, November 17–19, 2010, Revised Selected Papers, volume 6748 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22493-8 (print), 3-642-22494-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22494-2.

Doberkat:1981:SAS

[Dob81] Ernst-Erich Doberkat. Stochastic automata: stability, nondeterminism, and prediction, volume 113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10835-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 135 pp. LCCN QA267.5.P7 D63. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0113.htm;
http://www.springerlink.
com/content/978-0-387-
10835-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=113.

Dolev:2006:DCI

[Dol06] Shlomi Dolev, editor.

Distributed Computing: 20th International Symposium, DISC 2006, Stockholm, Sweden, September 18–20, 2006. Proceedings, volume 4167 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44624-9 (print), 3-540-44627-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
44627-9.

Dolev:2009:AAW

[Dol09] Shlomi Dolev, editor.

Algorithmic Aspects of Wireless Sensor Networks: 5th International Workshop, ALGOSEN-SORS 2009, Rhodes, Greece, July 10–11, 2009. Revised Selected Papers, volume 5804 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05433-1 (print), 3-642-05434-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
05434-1.

Donahue:1976:CDP

[Don76] James Edward Donahue. Complementary definitions of programming language semantics, volume 42 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1976. CODEN LNCSD9. ISBN 0-387-07628-X. ISSN 0302-9743 (print), 1611-3349 (electronic). 172 pp. LCCN QA76.7.D661. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0042.
htm;   http://www.springerlink.
com/content/978-0-387-
07628-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=42.

Donald:1989:EDR

[Don89] Bruce R. Donald. Error detection and recovery in robotics, volume 336 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-96909-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xxiv + 314 pp. LCCN TJ211.417 .D661 1989.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0336.
htm;   http://www.springerlink.
com/content/978-0-387-
96909-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=336.

Revised version of thesis (Ph. D.)–M.I.T., 1987.

Dorr:1995:EGR

[Dor95]  Heiko Dorr.  Effi-

cient graph rewriting and its implementation, volume 922 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 0-387-60055-8 (USA), 3-540-60055-8. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 266 pp. LCCN QA267.A1 L43 no.922. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0922.htm;
 http://www.springerlink.
com/content/978-0-387-
60055-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=922.

Dorronsoro:2002:ANN

[Dor02] Josè R. Dorronsoro, editor. Artificial Neural Networks—ICANN 2002: International Conference, Madrid, Spain, August 28–30, 2002. Proceedings, volume 2415 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44074-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I56 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2415.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44074-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2415.

Doshita:1993:ALT

[Dos93] S. Doshita, editor. Algorithmic learning theory: third workshop, ALT ’92, Tokyo, Japan, October 20–22, 1992: proceedings, volume 743 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57369-0 (Berlin), 0-387-57369-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A48 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0743.htm;
 http://www.springerlink.
com/content/978-3-540-
57369-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=743.

Diaz:1999:IDM

[DOS99] Michel Diaz, Philippe Owezarski, and P. (Patrick) Senac, editors. Interactive distributed multimedia systems and telecommunication services: 6th International Workshop, IDMS ’99, Toulouse, France, October 12–15, 1999: Proceedings, volume 1718 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66595-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I59 I586 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1718.
htm;   http://www.springerlink.
com/content/978-3-540-
66595-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1718.

Domaratzki:2005:IAA

[DOSY05]  Michael  Domaratzki,

Alexander Okhotin, Kai Salomaa, and Sheng Yu, editors. Implementation and application of automata: 9th international conference, CIAA 2004, Kingston, Canada, July 22–24, 2004: revised selected papers, volume 3317 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24318-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .C5242 2004. URL

                                        http://springerlink.
                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3317;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        24318-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3317.

Dovrolis:2005:PAN

[Dov05] Constantinos Dovrolis, editor. Passive and active network measurement: 6th international workshop, PAM 2005, Boston, MA, USA, March 31–April 1, 2005: Proceedings, volume 3431 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25520-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .P36 2005. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3431;   http:
//www.springerlink.
com/content/978-3-540-
25520-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3431;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b135479.

Dongarra:1997:VPP

[DP97] J. J. Dongarra and Jose M. L. M. Palma, editors. Vector and parallel processing—VECPAR ’96: Second International Conference on Vector and Parallel Processing-Systems and Applications, Porto, Portugal, September 25–27, 1996: Selected Papers, volume 1215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62828-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I552 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1215.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62828-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1215.

Dubois:2006:AIS

[DP06] Eric Dubois and Klaus Pohl, editors. Advanced Information Systems Engineering: 18th International Conference, CAiSE 2006, Luxembourg, Luxembourg, June 5–9, 2006. Proceedings, volume 4001 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34652-X (print), 3-540-34653-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34653-1.

DiTosto:2010:MAB

[DP10a] Gennaro Di Tosto and H. Van Dyke Parunak, editors. Multi-Agent-Based Simulation X: International Workshop, MABS 2009, Budapest, Hungary, May 11–12, 2009 Revised Selected Papers, volume 5683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13552-8 (print), 3-642-13553-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13553-8.

Dovier:2010:YPL

[DP10b] Agostino Dovier and Enrico Pontelli, editors. A 25–Year Perspective on Logic Programming: Achievements of the Italian Association for Logic Programming, GULP, volume 6125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14308-3 (print), 3-642-14309-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14309-0.

Delzanno:2011:RPI

[DP11] Giorgio Delzanno and Igor Potapov, editors. Reachability Problems: 5th International Workshop, RP 2011, Genoa, Italy, September 28– 30, 2011. Proceedings, volume 6945 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24287-1 (print), 3-642-24288-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24288-5.

Donaldson:2012:MCS

[DP12] Alastair Donaldson and David Parker, editors. Model Checking Software: 19th International Workshop, SPIN 2012, Oxford, UK, July 23– 24, 2012. Proceedings, volume 7385 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31758-8 (print), 3-642-31759-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31759-0.

Diaz:2005:CAS

[DPA05] Roberto Moreno Díaz, Franz Pichler, and Alexis Quesada Arencibia, editors. Computer aided systems theory — EU-ROCAST 2005: 10th international conference on computer aided systems theory, Las Palmas de Gran Canaria, Spain, February 7–11, 2005, revised selected papers, volume 3643 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29002-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29002-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3643.

Diaz:2007:CAS

[DPA07] Roberto Moreno Díaz, Franz Pichler, and Alexis Quesada Arencibia, editors. Computer Aided Systems Theory — EURO-CAST 2007: 11th International Conference on Computer Aided Systems Theory, Las Palmas de Gran Canaria, Spain, February 12–16, 2007, Revised Selected Papers, volume 4739 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75866-6 (print), 3-540-75867-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75867-9.

Deb:2004:GECa

[DPB^{+}04a] Kalyanmoy Deb, Riccardo Poli, Wolfgang Banzhaf, Hans-Georg Beyer, Edmund Burke, Paul Darwen, Dipankar Dasgupta, Dario Floreano, James Foster, Mark Harman, Owen Holland, Pier Luca Lanzi, Lee Spector, Andrea Tettamanzi, Dirk Thierens, and Andy Tyrrell, editors. Genetic and Evolutionary Computation — GECCO 2004: Genetic and Evolutionary Computation Conference, Seattle, WA, USA, June 26–30, 2004. Proceedings, Part I, volume 3102 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22344-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.623 .G45 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/

0558/tocs/t3102.htm;
 http://www.springerlink.
com/content/978-3-540-
22344-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3102;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98643.

Deb:2004:GECb

[DPB^{+}04b] Kalyanmoy Deb, Riccardo Poli, Wolfgang Banzhaf, Hans-Georg Beyer, Edmund Burke, Paul Darwen, Dipankar Dasgupta, Dario Floreano, James Foster, Mark Harman, Owen Holland, Pier Luca Lanzi, Lee Spector, Andrea Tettamanzi, Dirk Thierens, and Andy Tyrrell, editors. Genetic and Evolutionary Computation: GECCO 2004: Genetic and Evolutionary Computation Conference, Seattle, WA, USA, June 26–30, 2004. Proceedings, Part II, volume 3103 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22343-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.623 .G45 2004. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t3103.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22343-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3103;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98645.

Dayde:2007:HPC

[DPC^{+}07] Michel Daydé, José M. L. M. Palma, Álvaro L. G. A. Coutinho, Esther Pacitti, and João Correia Lopes, editors. High Performance Computing for Computational Science — VECPAR 2006: 7th International Conference, Rio de Janeiro, Brazil, June 10–13, 2006, Revised Selected and Invited Papers, volume 4395 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71350-6 (print), 3-540-71351-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71351-7.

Das:2008:NAH

[DPLW08] Amitabha Das, Hung Keng Pung, Francis Bu Sung Lee, and Lawrence Wai Choong Wong, editors. NET-WORKING 2008 Ad Hoc and Sensor Networks, Wireless Networks, Next Generation Internet: 7th International IFIP-TC6 Networking Conference Singapore, May 5–9, 2008 Proceedings, volume 4982 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79548-0 (print), 3-540-79549-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79549-0.

delPobil:1998:IEA

[dPMA98] Angel Pasqual del Pobil, Jose Mira, and Moonis Ali, editors. Industrial and engineering applications of artificial intelligence and expert systems: 11th International Conference IEA/AIE-98, Benicassim, Castellon, Spain, June 1998: Proceedings. Part II. Tasks and Methods in Applied Artificial Intelligence, volume 1416 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64574-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1415-1416. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1415.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64574-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1416.

Contents: v. 1. Methodology and tools in knowledge-based systems / Jose Mira, Angel Pasqual del Pobil, Moonis Ali, eds. — v. 2. Tasks and methods in applied artificial intelligence / Angel Pasqual del Pobil, Jose Mira, Moonis Ali, eds.

Domingo-Pascual:2011:NIIa

[DPMP^{+}11a] Jordi Domingo-Pascual, Pietro Manzoni, Sergio Palazzo, Ana Pont, and Caterina Scoglio, editors. NETWORKING 2011: 10th International IFIP TC 6 Networking Conference, Valencia, Spain, May 9–13, 2011, Proceedings, Part I, volume 6640 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20756-1 (print), 3-642-20757-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20757-0.

Domingo-Pascual:2011:NIIb

[DPMP^{+}11b] Jordi Domingo-Pascual, Pietro Manzoni, Sergio Palazzo, Ana Pont, and Caterina Scoglio, editors. NETWORKING 2011: 10th International IFIP TC 6 Networking Conference, Valencia, Spain, May 9–13, 2011, Proceedings, Part II, volume 6641 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20797-9 (print), 3-642-20798-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20798-3.

Donnelly:2012:IAS

[DPNM12] Mark Donnelly, Cristiano Paggetti, Chris Nugent, and Mounir Mokhtari, editors. Impact Analysis of Solutions for Chronic Disease Prevention and Management: 10th International Conference on Smart Homes and Health Telematics, ICOST 2012, Artiminio, Italy, June 12–15, 2012. Proceedings, volume 7251 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30778-7 (print), 3-642-30779-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30779-9.

Dix:1995:NME

[DPP95] J. Dix, Luis Moniz Pereira, and Teodor C. Przymusinski, editors. Non-monotonic extensions of logic programming: ICLP ’94 workshop, Santa Margherita Ligure, Italy, June 17, 1994: Selected Papers, volume 927 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59467-1 (Berlin), 0-387-59467-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.63 .N68 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0927.htm;
 http://www.springerlink.
com/content/978-3-540-
59467-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=927.

Dix:1997:NME

[DPP97] J. Dix, Luis Moniz Pereira, and Teodor C. Przymusinski, editors. Non-monotonic extensions of logic programming: 2nd International Workshop, NMELP ’96, Bad Honnef, Germany, September 5–6, 1996: Selected Papers, volume 1216 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62843-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .N65 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1216.htm;
 http://www.springerlink.
com/content/978-3-540-
62843-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1216.

Dix:1998:LPK

[DPP98] J. Dix, Luis Moniz Pereira, and Teodor C. Przymusinski, editors. Logic programming and knowledge representation: Third International Workshop, LPKR’97: Port Jefferson, New York, USA, October 17, 1997: Selected Papers, volume 1471 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64958-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1471. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1471.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64958-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1471.

DiGesu:2009:FLA

[DPP09] Vito Di Gesù, Sankar Kumar Pal, and Alfredo Petrosino, editors. Fuzzy Logic and Applications: 8th International Workshop, WILF 2009 Palermo, Italy, June 9–12, 2009 Proceedings, volume 5571 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02281-2 (print), 3-642-02282-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02282-1.

delPobil:1995:SRM

[dPS95] Angel Pasqual del Pobil and Miguel Angel Serna. Spatial representation and motion planning, volume 1014 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60620-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 242 pp. LCCN TJ211.4 .S63 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1014.htm;
http://www.springerlink.
com/content/978-3-540-
60620-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1014.

Domingo-Pascual:2011:TMA

[DPSU11] Jordi Domingo-Pascual, Yuval Shavitt, and Steve Uhlig, editors. Traffic Monitoring and Analysis: Third International Workshop, TMA 2011, Vienna, Austria, April 27, 2011. Proceedings, volume 6613 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20304-3 (print), 3-642-20305-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20305-3.

Dochev:2008:AIM

[DPT08] Danail Dochev, Marco Pistore, and Paolo Traverso, editors. Artificial Intelligence: Methodology, Systems, and Applications: 13th International Conference, AIMSA 2008, Varna, Bulgaria, September 4– 6, 2008. Proceedings, volume 5253 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85775-3 (print), 3-540-85776-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
85776-1.

DiMartino:2012:WWG

[DPT12] Sergio Di Martino, Adriano Peron, and Taro Tezuka, editors. Web and Wireless Geographical Information Systems: 11th International Symposium, W2GIS 2012, Naples, Italy, April 12– 13, 2012. Proceedings, volume 7236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29246-1 (print), 3-642-29247-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29247-7.

Defago:2011:SSS

[DPV11] Xavier Défago, Franck Petit, and Vincent Villain, editors. Stabilization, Safety, and Security of Distributed Systems: 13th International Symposium, SSS 2011, Grenoble, France, October 10–12, 2011. Proceedings, volume 6976 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24549-8 (print), 3-642-24550-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24550-3.

Desel:2004:BPM

[DPW04] Jörg Desel, Barbara Pernici, and Mathias Weske, editors. Business Process Management: Second International Conference, BPM 2004, Potsdam, Germany, June 17–18, 2004: Proceedings, volume 3080 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22235-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.2 .B674 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3080.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22235-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3080;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98280.

DePaoli:2012:SOC

[DPZ12] Flavio De Paoli, Ernesto Pimentel, and Gianluigi Zavattaro, editors. Service-Oriented and Cloud Computing: First European Conference, ESOCC 2012, Bertinoro, Italy, September 19–21, 2012. Proceedings, volume 7592 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33426-1 (print), 3-642-33427-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33427-6.

Deng:2002:ICS

[DQBZ02] Robert Deng, Sihan Qing, Feng Bao, and Jianying Zhou, editors. Information and Communications Security: 4th International Conference, ICICS 2002, Singapore, December 9– 12, 2002. Proceedings, volume 2513 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00164-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 .I554 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2513.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00164-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2513.

Diaz:1981:FPC

[DR81] J. Díaz and I. Ramos, editors. Formalization of programming concepts: international colloquium, Peniscola, Spain, April 19–25, 1981: proceedings, volume 107 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10699-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .F66. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0107.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        10699-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=107.

Daemen:2002:FSE

[DR02a] Joan Daemen and Vincent Rijmen, editors. Fast Software Encryption: 9th International Workshop, FSE 2002, Leuven, Belgium, February 4–6, 2002. Revised Papers, volume 2365 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44009-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F77 2002b. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2365.htm;
 http://www.springerlink.
com/content/978-3-540-
44009-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2365.

Diks:2002:MFC

[DR02b] Krzysztof Diks and Wojciech Rytter, editors. Mathematical Foundations of Computer Science 2002: 27th International Symposium, MFCS 2002, Warsaw, Poland, August 26– 30, 2002. Proceedings, volume 2420 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44040-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA267 .M38 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2420.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44040-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2420.

DeFelice:2005:DLT

[DR05] Clelia De Felice and Antonio Restivo, editors. Developments in language theory: 9th international conference, DLT 2005, Palermo, Italy, July 4–8, 2005: proceedings, volume 3572 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26546-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .C63 2005; QA267.3; QA267.3 .C63 2005eb; Internet. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26546-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3572.

DiCrescenzo:2006:FCD

[DR06] Giovanni Di Crescenzo and Avi Rubin, editors. Financial Cryptography and Data Security:

10th International Conference, FC 2006 Anguilla, British West Indies, February 27–March 2, 2006 Revised Selected Papers, volume 4107 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46255-4 (print), 3-540-46256-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46256-9.

DiNitto:2009:SOC

[DR09] Elisabetta Di Nitto and Matei Ripeanu, editors. Service-Oriented Computing — ICSOC 2007 Workshops: IC-SOC 2007, International Workshops, Vienna, Austria, September 17, 2007, Revised Selected Papers, volume 4907 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-93850-8 (print), 3-540-93851-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        93851-4.

DeMeuter:2011:CML

[DR11] Wolfgang De Meuter and Gruia-Catalin Roman, editors. Coordination Models and Languages: 13th International Conference, COORDINA-TION 2011, Reykjavik, Iceland, June 6–9, 2011. Proceedings, volume 6721 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21463-0 (print), 3-642-21464-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21464-6.

Costa:2011:AAI

[dRCVT11] Antônio Carlos da Rocha Costa, Rosa Maria Vicari, and Flavio Tonidandel, editors. Advances in Artificial Intelligence — SBIA 2010: 20th Brazilian Symposium on Artificial Intelligence, São Bernardo do Campo, Brazil, October 23–28, 2010. Proceedings, volume 6404 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16137-5 (print), 3-642-16138-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16138-4.

Debled-Rennesson:2011:DGC

[DRDKE11] Isabelle Debled-Rennesson, Eric Domenjoud, Bertrand Kerautret, and Philippe Even, editors. Discrete Geometry for Computer Imagery: 16th IAPR International Conference, DGCI 2011, Nancy, France, April 6–8, 2011. Proceedings, volume 6607 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19866-X (print), 3-642-19867-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19867-0.

deRaedt:2001:MLE

[dRF01] Luc de Raedt and Peter Flach, editors. Machine learning: ECML 2001: 12th European Conference on Machine Learning, Freiburg, Germany, September 5–7, 2001: Proceedings, volume 2167 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42536-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 2001; QA267.A1 L43 no.2167. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2167.htm;
 http://www.springerlink.
com/content/978-3-540-
42536-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2167.

deRoever:1998:CSD

[dRLP98] Willem Paul de Roever, Hans Langmaack, and Amir Pnueli, editors. Compositionality: The Significant Difference: International Symposium, COMPOS ’97, Bad Malente, Germany, September 8–12, 1997: revised lectures, volume 1536 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65493-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.C639 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1536.htm;
http://www.springerlink.
com/content/978-3-540-
65493-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1536.

Drossopoulou:2008:PLS

[Dro08] Sophia Drossopoulou, editor. Programming Languages and Systems: 17th European Symposium on Programming, ESOP 2008, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2008, Budapest, Hungary, March 29–April 6, 2008. Proceedings, volume 4960 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78738-0 (print), 3-540-78739-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78739-6.

Drossopoulou:2009:EOO

[Dro09] Sophia Drossopoulou, editor. ECOOP 2009

— Object-Oriented Programming: 23rd European Conference, Genoa, Italy, July 6–10, 2009. Proceedings, volume 5653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03012-2 (print), 3-642-03013-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03013-0.

Desel:2004:LCP

[DRR04] Jörg Desel, Wolfgang Reisig, and Grzegorz Rozenberg, editors. Lectures on Concurrency and Petri Nets: advances in Petri nets, volume 3098 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22261-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.L38 2004. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3098.
htm;   http://www.springerlink.
com/content/978-3-540-
22261-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3098;
 http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b98282.

deRaedt:2001:PDM

[dRS01] Luc de Raedt and Arno Siebes, editors. Principles of data mining and knowledge discovery: 5th European Conference, PKDD 2001, Freiburg, Germany, September 3– 5, 2001: Proceedings, volume 2168 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42534-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 P16 2001; QA267.A1 L43 no.2168. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2168.htm;
http://www.springerlink.
com/content/978-3-540-
42534-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2168.

Dobbertin:2005:AES

[DRS05] Hans Dobbertin, Vincent Rijmen, and Aleksandra Sowa, editors. Advanced encryption standard — AES: 4th international conference, AES 2004, Bonn, Germany, May 10–12, 2004, revised selected and invited papers, volume 3373 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26557-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26557-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3373.

Dumas:2008:BPM

[DRS08] Marlon Dumas, Manfred Reichert, and Ming-Chien Shan, editors. Business Process Management: 6th International Conference, BPM 2008, Milan, Italy, September 2–4, 2008. Proceedings, volume 5240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85757-5 (print), 3-540-85758-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85758-7.

Dumke:2001:PES

[DRSS01] Reiner Dumke, Claus Rautenstrauch, Andreas Schmietendorf, and André Scholz, editors. Performance Engineering: State of the Art and Current Trends, volume 2047 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42145-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 348 pp. LCCN QA76.9.E94 P425 2001; QA267.A1 L43 no.2047. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2047.htm;
http://www.springerlink.
com/content/978-3-540-
42145-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2047.

deRuyter:2010:AIF

[dRWK^{+}10] Boris de Ruyter, Reiner Wichert, David V. Keyson, Panos Markopoulos, Norbert Streitz, Monica Divitini, Nikolaos Georgantas, and Antonio Mana Gomez, editors. Ambient Intelligence: First International Joint Conference, AmI 2010, Malaga, Spain, November 10– 12, 2010. Proceedings, volume 6439 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16916-3 (print), 3-642-16917-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16917-5.

Diaz:1996:AEF

[DS96] J. (Josep) Diaz and Maria Serna, editors. Algorithms — ESA ’96: Fourth Annual European Symposium, Barcelona, Spain, September 25– 27, 1996: proceedings, volume 1136 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61680-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1136.htm;
 http://www.springerlink.
com/content/978-3-540-
61680-1;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1136.

Desel:1998:ATP

[DS98] Jörg Desel and Manuel Silva, editors. Application and theory of Petri nets 1998: 19th International Conference, ICATPN ’98, Lisbon, Portugal, June 22– 26, 1998: Proceedings, volume 1420 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64677-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1420.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1420.
htm;   http://www.springerlink.
com/content/978-3-540-
64677-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1420.

Delugach:2001:CSB

[DS01a] Harry S. Delugach and Gerd Stumme, editors. Conceptual structures: broadening the base: 9th International Conference on Conceptual Structures, ICCS 2001 Stanford, CA, USA, July 30– August 3, 2001: Proceedings, volume 2120 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42344-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2 .I58 2001; QA267.A1 L43 no.2120. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2120.htm;
 http://www.springerlink.
com/content/978-3-540-
42344-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2120.

Dignum:2001:AMEa

[DS01b] Frank Dignum and Carles Sierra. Agent mediated electronic commerce: the European AgentLink perspective, volume 1991 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41671-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 239 pp. LCCN QA76.76.I58 A32 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1991.

htm;   http://www.springerlink.
com/content/978-3-540-
41671-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1991.

deSwart:2002:RMC

[dS02] Harrie M. C. de Swart, editor. Relational Methods in Computer Science: 6th International Conference, RelMiCS 2001 and 1st Workshop of COST Action 274 TARSKI, Oisterwijk, The Netherlands, October 16– 21, 2001. Revised Papers, volume 2561 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00315-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .R437 2001; QA75.5 .R44 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2561.htm;
http://www.springerlink.
com/content/978-3-540-
00315-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2561.

Also available via the World Wide Web.

Dingledine:2003:PETa

[DS03] Roger Dingledine and Paul F. Syverson, editors. Privacy enhancing technologies: Second International Workshop, PET 2002, San Francisco, CA, USA, April 2002: Revised Papers, volume 2482 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00565-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 P443 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2482.htm;
 http://www.springerlink.
com/content/978-3-540-
00565-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2482.

Danos:2005:CMS

[DS05a] Vincent Danos and Vincent Schachter, editors. Computational methods in systems biology: International Conference CMSB 2004, Paris, France, May 26– 28, 2004: revised selected papers, volume 3082 of Lecture Notes in Bioinformatics and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25375-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH506 .C67 2004. URL http:

//www.springerlink.
com/content/978-3-540-
25375-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3082.

DeNicola:2005:TGC

[DS05b] Rocco De Nicola and Davide Sangiorgi, editors. Trustworthy global computing: international symposium, TGC 2005, Edinburgh, UK, April 7– 9, 2005: revised selected papers, volume 3705 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30007-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 T435 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30007-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3705.

Dunne:2006:UTP

[DS06] Steve Dunne and Bill Stoddart, editors. Unifying Theories of Programming: First International Symposium, UTP 2006, Walworth Castle, County Durham, UK, February 5–7, 2006, Revised Selected Papers, volume 4010 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34750-X (print), 3-540-34752-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34752-1.

Dzeroski:2007:KDI

[DS07] Sašo Džeroski and Jan Struyf, editors. Knowledge Discovery in Inductive Databases: 5th International Workshop, KDID 2006 Berlin, Germany, September 18, 2006 Revised Selected and Invited Papers, volume 4747 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75548-9 (print), 3-540-75549-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75549-4.

Dillenbourg:2008:TCT

[DS08] Pierre Dillenbourg and Marcus Specht, editors. Times of Convergence. Technologies Across Learning Contexts: Third European Conference on Technology Enhanced Learning, EC-TEL 2008, Maastricht, The Netherlands, September 16–19, 2008. Proceedings, volume 5192 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87604-9 (print), 3-540-87605-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87605-2.

David:2009:MAB

[DS09a] Nuno David and Jaime Simão Sichman, editors. Multi-Agent-Based Simulation IX: International Workshop, MABS 2008, Estoril, Portugal, May 12–13, 2008, Revised Selected Papers, volume 5269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01990-0 (print), 3-642-01991-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01991-3.

Deaton:2009:DCM

[DS09b] Russell Deaton and Akira Suyama, editors. DNA Computing and Molecular Programming: 15th International Conference, DNA 15, Fayetteville, AR, USA, June 8–11, 2009, Revised Selected Papers, volume 5877 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10603-X (print), 3-642-10604-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10604-0.

deSwart:2011:RAM

[dS11a] Harrie de Swart, editor. Relational and Algebraic Methods in Computer Science: 12th International Conference, RAMICS 2011, Rotterdam, The Netherlands, May 30 — June 3, 2011. Proceedings, volume 6663 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21069-4 (print), 3-642-21070-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21070-9.

Dingel:2011:MSE

[DS11b] Juergen Dingel and Arnor Solberg, editors. Models in Software Engineering: Workshops and Symposia at MODELS 2010, Oslo, Norway, October 2–8, 2010, Reports and Revised Selected Papers, volume 6627 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21209-3 (print), 3-642-21210-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21210-9.

Domaratzki:2011:IAA

[DS11c] Michael Domaratzki and Kai Salomaa, editors. Implementation and Application of Automata: 15th International Conference, CIAA 2010, Winnipeg, MB, Canada, August 12–15, 2010. Revised Selected Papers, volume 6482 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18097-3 (print), 3-642-18098-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18098-9.

deSa:2007:ANNa

[dSADM07a] Joaquim Marques de Sá, Luís A. Alexandre, Włodzisław Duch, and Danilo Mandic, editors. Artificial Neural Networks — ICANN 2007: 17th International Conference, Porto, Portugal, September 9–13, 2007, Proceedings, Part I, volume 4668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74689-7 (print), 3-540-74690-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74690-4.

deSa:2007:ANNb

[dSADM07b] Joaquim Marques de Sá, Luís A. Alexandre, Włodzisław Duch, and Danilo Mandic, editors. Artificial Neural Networks — ICANN 2007: 17th International Conference, Porto, Portugal, September 9–13, 2007, Proceedings, Part II, volume 4669 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74693-5 (print), 3-540-74695-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74695-9.

Davies:2004:FMS

[DSB04] Jim Davies, Wolfram Schulte, and Mike Barnett, editors. Formal Methods and Software Engineering: 6th International Conference on Formal Engineering Methods, ICFEM 2004, Seattle, WA, USA, November 8–12, 2004. Proceedings, volume 3308 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23841-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3308.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23841-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3308;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b102837.

DeDecker:2010:CMS

[DSB10] Bart De Decker and Ingrid Schaumüller-Bichl, editors. Communications and Multimedia Security: 11th IFIP TC 6/TC 11 International Conference, CMS 2010, Linz, Austria, May 31

— June 2, 2010. Proceedings, volume 6109 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13240-5 (print), 3-642-13241-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13241-4.

Ruiz-del-Solar:2011:RRS

[dSCP11] Javier Ruiz del Solar, Eric Chown, and Paul G. Plöger, editors. RoboCup 2010:

Robot Soccer World Cup XIV, volume 6556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20216-0 (print), 3-642-20217-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20217-9.

deSouza:2004:TNI

[dSDL04] José Neuman de Souza, Petre Dini, and Pascal Lorenz, editors. Telecommunications and Networking—ICT 2004: 11th International Conference on Telecommunications, Fortaleza, Brazil, August 1–6, 2004: Proceedings, volume 3124 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22571-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5101.A1 I482975 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3124.
htm;   http://www.springerlink.
com/content/978-3-540-
22571-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=3124;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99377.

Das:2004:DCI

[DSDS04] Nabanita Das, Arunabha Sen, Sajal K. Das, and Bhabani P. Sinha, editors. Distributed Computing — IWDC 2004: 6th International Workshop, Kolkata, India, December 27–30, 2004: Proceedings, volume 3326 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24076-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I68 2004.

                                        URL http://springerlink.
                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3326;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        24076-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3326.

deCapitanidiVimercati:2005:CSE

[dSG05] Sabrina de Capitani di Vimercati, Paul Syverson, and Dieter Gollmann, editors. Computer security — ESORICS 2005: 10th European symposium on research in computer security, Milan, Italy, September 12–14, 2005: proceedings, volume 3679 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28963-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28963-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3679.

Dastani:2011:LMD

[DSHL11] Mehdi Dastani, Amal El Fallah Seghrouchni, Jomi Hübner, and João Leite, editors. Languages, Methodologies, and Development Tools for Multi-Agent Systems: Third International Workshop, LADS 2010, Lyon, France, August 30 — September 1, 2010, Revised Selected Papers, volume 6822 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22722-8 (print), 3-642-22723-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22723-3.

deSouto:2012:ABC

[dSK12] Marcilio C. de Souto and Maricel G. Kann, editors. Advances in Bioinformatics and Computational Biology: 7th Brazilian Symposium on Bioinformatics, BSB 2012, Campo Grande, Brazil, August 15–17, 2012. Proceedings, volume 7409 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31926-2 (print), 3-642-31927-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31927-3.

Dohi:2008:MIA

[DSL08] Takeyoshi Dohi, Ichiro Sakuma, and Hongen Liao, editors. Medical Imaging and Augmented Reality: 4th International Workshop Tokyo, Japan, August 1–2, 2008 Proceedings, volume 5128 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79981-8 (print), 3-540-79982-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79982-5.

Dastani:2008:LMD

[DSLT08] Mehdi Dastani, Amal El Fallah Seghrouchni, João Leite, and Paolo Torroni, editors. Languages, Methodologies and Development Tools for Multi-Agent Systems: First International Workshop, LADS 2007, Durham, UK, September 4–6, 2007. Revised Selected Papers, volume 5118 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85057-0 (print), 3-540-85058-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85058-8.

Dastani:2010:LMD

[DSLT10] Mehdi Dastani, Amal El Fallah Seghrouchni, João Leite, and Paolo Torroni, editors. Languages, Methodologies, and Development Tools for Multi-Agent Systems: Second International Workshop, LADS 2009, Torino, Italy, September 7–9, 2009, Revised Selected Papers, volume 6039 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13337-1 (print), 3-642-13338-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13338-1.

Davidor:1994:PPS

[DSM94] Yuval Davidor, Hans-Paul Schwefel, and Reinhard Manner, editors. Parallel problem solving from nature — PPSN III: International Conference on Evolutionary Computation, the Third Conference on Parallel Problem Solving from Nature, Jerusalem, Israel, October 9–14, 1994: proceedings, volume 866 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58484-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I535 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0866.htm;
 http://www.springerlink.
com/content/978-3-540-
58484-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=866.

Dey:2003:UUC

[DSM03] Anind K. Dey, Albrecht Schmidt, and Joseph F. McCarthy, editors. Ubi-Comp 2003: Ubiquitous Computing: 5th International Conference, Seattle, WA, USA, October 12–15, 2003: Proceedings, volume 2864 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20301-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5915 .U25 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2864.
htm;   http://www.springerlink.
com/content/978-3-540-
20301-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2864;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b93949.

deSwart:2004:TAR

[dSOSR04] Harrie de Swart, Ewa Orlowska, Gunther Schmidt, and Marc Roubens, editors. Theory and Applications of Relational Structures as Knowledge Instruments: COST Action 274, TARSKI: Revised Papers, volume 2929 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20780-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .T483 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2929.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20780-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2929;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b94817.

deSwart:2006:TAR

[dSOSR06] Harrie de Swart, Ewa Orłowska, Gunther Schmidt, and Marc Roubens, editors. Theory and Applications of Relational Structures as Knowledge Instruments II: International Workshops of COST Action 274, TARSKI, 2002-2005, Selected Revised Papers, volume 4342 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69223-1 (print), 3-540-69224-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69224-9.

Dastani:2008:PMA

[DSRW08] Mehdi Dastani, Amal El Fallah Seghrouchni, Alessandro Ricci, and Michael Winikoff, editors. Programming Multi-Agent Systems: 5th International Workshop, ProMAS 2007 Honolulu, HI, USA, May 15, 2007 Revised and Invited Papers, volume 4908 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79042-X (print), 3-540-79043-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79043-3.

Dehne:1989:ADS

[DSS89] F. Dehne, J.-R. Sack, and N. Santoro, editors. Algorithms and data structures: Workshop WADS ’89, Ottawa, Canada, August 17–19, 1989: proceedings, volume 382 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51542-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 W68 1989. DM79.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0382.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51542-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=382.

Dehne:1991:ADS

[DSS91] F. Dehne, N. Santoro, and J.-R. Sack, editors. Algorithms and data structures: 2nd workshop, WADS ’91, Ottawa, Canada, August 14–16, 1991: proceedings, volume 519 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54343-0 (Berlin), 0-387-54343-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 A43 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0519.htm;
 http://www.springerlink.
com/content/978-3-540-
54343-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=519.

Dehne:2003:ADS

[DSS03] Frank Dehne, Jörg Rüdiger Sack, and Michiel Smid, editors. Algorithms and Data Structures: 8th International Workshop, WADS 2003, Ottawa, Ontario, Canada, July 30–August 1, 2003: Proceedings, volume 2748 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40545-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 W333 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2748.
htm;   http://www.springerlink.
com/content/978-3-540-
40545-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2748;

                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11837.

Dubitzky:2007:DHP

[DSS^{+}07] Werner Dubitzky, Assaf Schuster, Peter M. A. Sloot, Michael Schroeder, and Mathilde Romberg, editors. Distributed, High-Performance and Grid Computing in Computational Biology: International Workshop, GCCB 2006, Eilat, Israel, January 21, 2007. Proceedings, volume 4360 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69841-8 (print), 3-540-69968-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69968-2.

Davies:2011:FMF

[DSS11] Jim Davies, Leila Silva, and Adenilso Simao, editors. Formal Methods: Foundations and Applications: 13th Brazilian Symposium on Formal Methods, SBMF 2010, Natal, Brazil, November 8–11, 2010, Revised Selected Papers, volume 6527 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19828-7 (print), 3-642-19829-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19829-8.

Denko:2007:EDE

[DsSL^{+}07] Mieso K. Denko, Chi sheng Shih, Kuan-Ching Li, Shiao-Li Tsao, Qing-An Zeng, Soo Hyun Park, Young-Bae Ko, Shih-Hao Hung, and Jong Hyuk Park, editors. Emerging Directions in Embedded and Ubiquitous Computing: EUC 2007 Workshops: TRUST, WSOC, NCUS, UUWSN, USN, ESO, and SECUBIQ, Taipei, Taiwan, December 17– 20, 2007. Proceedings, volume 4809 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77089-5 (print), 3-540-77090-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        77090-9.

Dehne:2001:ADS

[DST01] F. Dehne, J.-R. Sack, and Roberto Tamassia, editors. Algorithms and Data Structures: 7th International Workshop, WADS 2001, Providence, RI, USA, August, 8–10, 2001, Proceedings, volume 2125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42423-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 W33 2001; QA267.A1 L43 no.2125.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2125.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        42423-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2125.

deSouza:2011:ABC

[dSTP11] Osmar Norberto de Souza, Guilherme P. Telles, and Mathew Palakal, editors. Advances in Bioinformatics and Computational Biology: 6th Brazilian Symposium on Bioinformatics, BSB 2011, Brasilia, Brazil, August 10–12, 2011. Proceedings, volume 6832 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22824-0 (print), 3-642-22825-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22825-4.

DiBenedetto:2001:HSC

[DSV01] Maria Domenica Di Benedetto and Alberto Sangiovanni-Vincentelli, editors. Hybrid systems: computation and control: 4th International Workshop, HSCC 2001, Rome, Italy, March 28– 30, 2001: Proceedings, volume 2034 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41866-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38 .H73 2001; QA267.A1 L43 no.2034. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2034.htm;
 http://www.springerlink.
com/content/978-3-540-
41866-5;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2034.

Dai:2004:AKD

[DSZ04] Honghua Dai, Ramakrishnan Srikant, and Chengqi Zhang, editors. Advances in Knowledge Discovery and Data Mining: 8th Pacific-Asia Conference, PAKDD 2004, Sydney, Australia, May 26–28, 2004: Proceedings, volume 3056 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22064-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3056.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22064-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3056;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97861.

Dehne:2007:ADS

[DSZ07] Frank Dehne, Jörg-Rüdiger Sack, and Norbert Zeh, editors. Algorithms and Data Structures. 10th International Workshop, WADS 2007, Halifax, Canada, August 15–17, 2007. Proceedings, volume 4619. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73948-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73948-7.

Ding:2010:NPC

[DSZ10] Chen Ding, Zhiyuan Shao, and Ran Zheng, editors. Network and Parallel Computing: IFIP International Conference, NPC 2010, Zhengzhou, China, September 13– 15, 2010. Proceedings, volume 6289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15671-1 (print), 3-642-15672-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15672-4.

Datta:2011:SIT

[DSZ^{+}11] Anwitaman Datta, Stuart Shulman, Baihua Zheng, Shou-De Lin, Aixin Sun, and Ee-Peng Lim, editors. Social Informatics: Third International Conference, SocInfo 2011, Singapore, October 6–8, 2011. Proceedings, volume 6984 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24703-2 (print), 3-642-24704-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24704-0.

Demetrovics:1989:MSM

[DT89] Janos Demetrovics and B. Thalheim, editors. MFDBS 89: 2nd Symposium on Mathematical Fundamentals of Database Systems, Visegrad, Hungary, June 26– 30, 1989: proceedings, volume 364 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51251-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S975 1989. DM57.50. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0364.htm;
 http://www.springerlink.
com/content/978-0-387-

51251-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=364.

Donatelli:2006:PNO

[DT06a] Susanna Donatelli and P. S. Thiagarajan, editors. Petri Nets and Other Models of Concurrency — ICATPN 2006: 27th International Conference on Applications and Theory of Petri Nets and Other Models of Concurrency, Turku, Finland, June 26–30, 2006. Proceedings, volume 4024 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34699-6 (print), 3-540-34700-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34700-2.

Durand:2006:SAS

[DT06b] Bruno Durand and Wolfgang Thomas, editors. STACS 2006: 23rd Annual Symposium on Theoretical Aspects of Computer Science, Marseille, France, February 23– 25, 2006. Proceedings, volume 3884 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32301-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32301-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3884.

Dzeroski:2007:CDS

[DT07] Sašo Džeroski and Ljupčo Todorovski, editors. Computational Discovery of Scientific Knowledge: Introduction, Techniques, and Applications in Environmental and Life Sciences, volume 4660 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73919-X (print), 3-540-73920-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73920-3.

Drogoul:1998:CRF

[DTF98] Alexis Drogoul, Milind Tambe, and T. Fukuda, editors. Collective robotics: First International Workshop, CRW ’98, Paris, France, July 4–5, 1998: Proceedings, volume 1456 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64768-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ210.3 .C79 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1456.htm;
http://www.springerlink.
com/content/978-3-540-
64768-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1456.

Dijkstra:2010:PRB

[DTMH10] Tjeerd M. H. Dijkstra, Evgeni Tsivtsivadze, Elena Marchiori, and Tom Heskes, editors. Pattern Recognition in Bioinformatics: 5th IAPR International Conference, PRIB 2010, Nijmegen, The Netherlands, September 22–24, 2010. Proceedings, volume 6282 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16000-X (print), 3-642-16001-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16001-1.

Deprettere:2002:EPD

[DTV02] Ed. F. Deprettere, Jurgen Teich, and Stamatis Vassiliadis, editors. Embedded processor design challenges: systems, architectures, modeling, and simulation — SAMOS, volume 2268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43322-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 D47 2002; TK7895.E42 S36 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2268.
htm;   http://www.springerlink.
com/content/978-3-540-
43322-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2268.

SAMOS 2001 took place from July 16–18, 2001 at the Research and Teaching Institute of East Aegean (INEAG) in Agios Konstantinos on the Island of Samos, Greece.

Dong:2003:AWA

[DTW03] Guozhu Dong, Chanjie Tang, and Wei Wang, editors. Advances in Web-Age Information Management: 4th International Conference, WAIM 2003, Chengdu, China, August 17–19, 2003: Proceedings, volume 2762 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40715-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2762.htm;
http://www.springerlink.
com/content/978-3-540-
40715-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2762;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11939.

Du:2000:CCA

[Du00] Dingzhu Du, editor.

Computing and combinatorics: 6th annual International Conference, COCOON 2000, Sydney, Australia, July 26– 28, 2000: Proceedings, volume 1858 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67787-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1858. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1858.htm;
 http://www.springerlink.
com/content/978-3-540-
67787-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1858.

Das:2009:WAC

[DU09] Sandip Das and Ryuhei Uehara, editors. WAL-COM: Algorithms and Computation: Third International Workshop, WALCOM 2009, Kolkata, India, February 18– 20, 2009. Proceedings, volume 5431 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00201-3 (print), 3-642-00202-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00202-1.

Dubois:2009:TPT

[Dub09] Catherine Dubois, editor. Tests and Proofs: Third International Conference, TAP 2009, Zurich, Switzerland, July 2–3, 2009. Proceedings, volume 5668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02948-5 (print), 3-642-02949-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02949-3.

Duesterwald:2004:CCI

[Due04] Evelyn Duesterwald, editor. Compiler Construction: 13th International Conference, CC 2004, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2004, Barcelona, Spain, March 29–April 2, 2004: Proceedings, volume 2985 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21297-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 C32 2004.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2985.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        21297-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2985;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95956.

Duffy:2007:DHM

[Duf07] Vincent G. Duffy, editor. Digital Human Modeling: First International Conference on Digital Human Modeling, ICDHM 2007, Held as Part of HCI International 2007, Beijing, China, July 22–27, 2007. Proceedings, volume 4561 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73318-3 (print), 3-540-73321-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73321-8.

Duffy:2009:DHM

[Duf09] Vincent G. Duffy, editor. Digital Human Modeling: Second International Conference, ICDHM 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19– 24, 2009. Proceedings, volume 5620 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02808-X (print), 3-642-02809-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02809-0.

Duffy:2011:DHM

[Duf11] Vincent G. Duffy, editor.

Digital Human Modeling: Third International Conference, ICDHM 2011, Held as Part of HCI International 2011, Orlando, FL, USA July 9– 14, 2011. Proceedings, volume 6777 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21798-2 (print), 3-642-21799-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21799-9.

Dunkelman:2009:FSE

[Dun09] Orr Dunkelman, editor. Fast Software Encryption: 16th International Workshop, FSE 2009 Leuven, Belgium, February 22–25, 2009 Revised Selected Papers, volume 5665 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03316-4 (print), 3-642-03317-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03317-9.

Dunkelman:2012:TCC

[Dun12] Orr Dunkelman, editor. Topics in Cryptology — CT-RSA 2012: The Cryptographers’ Track at the RSA Conference 2012, San Francisco, CA, USA, February 27 — March 2, 2012. Proceedings, volume 7178 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27953-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.

com/content/978-3-642-
27953-9.

Duran:2012:RLA

[Dur12] Franciso Durán, editor. Rewriting logic and its applications: 9th International Workshop, WRLA 2012, held as a satellite event of ETAPS, Tallinn, Estonia, March 24–25, 2012, Revised selected papers, volume 7571 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34004-0, 3-642-34005-9 (e-book). LCCN QA76.9.L63 W75 2012.

URL http://www.springerlink.
com/content/978-3-642-
34004-8;   http://www.
springerlink.com/content/
978-3-642-34005-5.

Dawson:2005:PCM

[DV05] Ed Dawson and Serge Vaudenay, editors. Progress in cryptology — MY-CRYPT 2005: first international conference on cryptology in Malaysia, Kuala Lumpur, Malaysia, September 28–30, 2005, proceedings, volume 3715 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28938-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28938-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3715.

Derrick:2007:FTN

[DV07a] John Derrick and Jüri Vain, editors. Formal Techniques for Networked and Distributed Systems — FORTE 2007: 27th IFIP WG 6.1 International Conference, Tallinn, Estonia, June 27–29, 2007. Proceedings, volume 4574 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73195-4 (print), 3-540-73196-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73196-2.

Dershowitz:2007:LPA

[DV07b] Nachum Dershowitz and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 14th International Conference, LPAR 2007, Yerevan, Armenia, October 15–19, 2007. Proceedings, volume 4790 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75558-6 (print), 3-540-75560-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75560-9.

Degano:2009:FASb

[DV09] Pierpaolo Degano and Luca Viganò, editors. Foundations and Applications of Security Analysis: Joint Workshop on Automated Reasoning for Security Protocol Analysis and Issues in the Theory of Security, ARSPA-WITS 2009, York, UK, March 28–29, 2009, Revised Selected Papers, volume 5511 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03458-6 (print), 3-642-03459-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03459-6.

Dawar:2010:CSL

[DV10] Anuj Dawar and Helmut Veith, editors. Computer Science Logic: 24th International Workshop, CSL 2010, 19th Annual Conference of the EACSL, Brno, Czech Republic, August 23– 27, 2010. Proceedings, volume 6247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15204-X (print), 3-642-15205-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15205-4.

Dignum:2006:ACI

[DvEF06] Frank P. M. Dignum, Rogier M. van Eijk, and Roberto Flores, editors. Agent Communication II: International Workshops on Agent Communication, AC 2005 and AC 2006, Utrecht, Netherlands, July 25, 2005, and Hakodate, Japan, May 9, 2006, Selected and Revised Papers, volume 3859 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68142-6 (print), 3-540-68143-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68143-4.

deVreede:2004:GDI

[dVGR04] Gert-Jan de Vreede, Luis A. Guerrero, and Gabriela Marin Raventos, editors. Groupware: Design, Implementation and Use: 10th International workshop, CRIWG 2004, San Carlos, Costa Rica, September 5–9, 2004: Proceedings, volume 3198 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23016-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 S356 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3198.
htm;   http://www.springerlink.
com/content/978-3-540-
23016-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3198;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100403.

Lobo:2008:SSS

[dVLKR^{+}08] Niels da Vitoria Lobo, Takis Kasparis, Fabio Roli, James T. Kwok, Michael Georgiopoulos, Georgios C. Anagnostopoulos, and Marco Loog, editors. Structural, Syntactic, and Statistical Pattern Recognition: Joint IAPR International Workshop, SSPR & SPR 2008, Orlando, USA, December 4–6, 2008. Proceedings, volume 5342 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89688-0 (print), 3-540-89689-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89689-0.

Dickmann:2011:SGI

[DVM^{+}11] Lutz Dickmann, Gerald Volkmann, Rainer Malaka, Susanne Boll, Antonio Krüger, and Patrick Olivier, editors. Smart Graphics: 11th International Symposium, SG 2011, Bremen, Germany, July 18–20, 2011. Proceedings, volume 6815 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22570-5 (print), 3-642-22571-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22571-0.

Dorst:1996:RUR

[DvV96] Leo Dorst, Michiel van Lambalgen, and Frans Voorbraak, editors. Reasoning with uncertainty in robotics: international workshop, RUR ’95, Amsterdam, The Netherlands, December 4–6, 1995: proceedings, volume 1093 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61376-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ210.3 .R87 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1093.htm;
http://www.springerlink.
com/content/978-3-540-
61376-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1093.

Contents: Mathematical foundations of navigation and perception for an autonomous mobile robot / J.L. Crowley — Reasoning with uncertainty in AI / F. Voorbraak — Robot navigation: integrating perception, environmental constraints and task execution within a probabilistic framework / A. Elfes — Uncertainty reasoning in object recognition by image processing / R. Hummel — Partially observable Markov decision processes for artificial intelligence / L.P. Kaelbling, M.L. Littman, A.R. Cassandra — An evidential approach to probabilistic map-building / D. Pagac, E.M. Nebot, H. Durrant-Whyte — Belief formation by constructing models / J. Paris, A. Vencovska — Causal relevance / G. Shafer — The robot control strategy in a domain with dynamical obstacles / G.V. Alferov, O.A. Malafeyev — Reasoning about noisy sensors (and effectors) in the situation calculus / F. Bacchus, J.Y. Halpern, H.J. Levesque — Recursive total least squares: an alternative to using the discrete Kalman filter in robot navigation / D.L. Boley, E.S. Steinmetz, K.T. Sutherland — A sensor-based motion planner for mobile robot navigation with uncertainty / B. Bouilly, T. Simeon — Knowledge considerations in robotics / R.I. Brafman, Y. Shoham — Neural network applications in sensor fusion for an autonomous mobile robot / J.W.M. van Dam, B.J.A. Krose, F.C.A. Groen — Structuring uncertain knowledge with hierarchical Bayesian networks / T. Frohlinghaus — Uncertainty treatment in a surface filling mobile robot / E. Gonzalez … [et al.] — Probabilistic map learning: necessity and difficulties / P. Hebert, S. Betge-Brezetz, R. Chatila — Robot navigation with Markov models: a framework for path planning and learning with limited computational resources / S. Koenig, R. Goodwin, R.G. Simmons — A refined method for occupancy grid interpretation / K. Konolige — Sensor planning with Bayesian decision theory / S. Kristensen — Perceptionbased self-localization using fuzzy locations / A. Saffiotti, L.P. Wesley.

Diekert:2007:CST

[DVV07] Volker Diekert, Mikhail V. Volkov, and Andrei Voronkov, editors. Computer Science — Theory and Applications: Second International Symposium on Computer Science in Russia, CSR 2007, Ekaterinburg, Russia, September 3–7, 2007. Proceedings, volume 4649 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74509-2 (print), 3-540-74510-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74510-5.

Darzentas:2008:AIT

[DVVA08] John Darzentas, George A. Vouros, Spyros Vosinakis, and Argyris Arnellos, editors. Artificial Intelligence: Theories, Models and Applications: 5th Hellenic Conference on AI, SETN 2008, Syros, Greece, October 2–4, 2008. Proceedings, volume 5138 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87880-7 (print), 3-540-87881-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87881-0.

Dongarra:1994:PSC

[DW94] J. J. Dongarra and Jerzy Waśniewski, editors. Parallel scientific computing: First International Workshop, PARA ’94, Lyngby, Denmark, June 20–23, 1994: proceedings, volume 879 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0879.
htm;   http://www.springerlink.
com/content/978-3-540-
58712-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=879.

Dahl:2003:PAD

[DW03a] Veronica Dahl and Philip Wadler, editors. Practical aspects of declarative languages: 5th International Symposium, PADL 2003, New Orleans, LA, USA, January 13–14, 2003: Proceedings, volume 2562 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00389-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .P32 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2562.htm;
 http://www.springerlink.
com/content/978-3-540-
00389-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2562.

Also available via the World Wide Web.

Dong:2003:FMS

[DW03b] Jin Song Dong and Jim Woodcock, editors. Formal Methods and Software Engineering: 5th International Conference on Formal Engineering Methods, ISFEM 2003, Singapore, November 5–7, 2003: Proceedings, volume 2885 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20461-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 I5788 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2885.
htm;   http://www.springerlink.
com/content/978-3-540-
20461-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2885;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94115.

Draheim:2006:TEA

[DW06] Dirk Draheim and Gerald Weber, editors. Trends in Enterprise Application Architecture: VLDB Workshop, TEAA 2005, Trondheim, Norway, August 28, 2005, Revised Selected Papers, volume 3888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32734-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32734-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3888.

Dawson:2007:ISP

[DW07a] Ed Dawson and Duncan S. Wong, editors. Information Security Practice and Experience: Third International Conference, ISPEC 2007, Hong Kong, China, May 7–9, 2007. Proceedings, volume 4464 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72159-2 (print), 3-540-72163-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72163-5.

Draheim:2007:TEA

[DW07b] Dirk Draheim and Gerald Weber, editors. Trends in Enterprise Application Architecture: 2nd International Conference, TEAA 2006, Berlin, Germany, November 29 — December 1, 2006, Revised Selected Papers, volume 4473 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75911-5 (print), 3-540-75912-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75912-6.

Dix:2010:MST

[DW10] Jürgen Dix and Cees Witteveen, editors. Multiagent System Technologies: 8th German Conference, MATES 2010, Leipzig, Germany, September 27–29, 2010. Proceedings, volume 6251 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16177-4 (print), 3-642-16178-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16178-0.

DeVos:2012:RCT

[DW12] Alexis De Vos and Robert Wille, editors. Reversible Computation: Third International Workshop, RC 2011, Gent, Belgium, July 4–5, 2011. Revised Papers, volume 7165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29516-9 (print), 3-642-29517-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29517-1.

Dwarkadas:2000:LCR

[Dwa00] Sandhya Dwarkadas, editor. Languages, Compilers, and Run-Time Systems for Scalable Computers: 5th International Workshop, LCR 2000, Rochester, NY, USA, May 25– 27, 2000. Selected Papers, volume 1915 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41185-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W658 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1915.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41185-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1915.

Desmedt:2005:CNS

[DWML05] Yvo G. Desmedt, Huaxiong Wang, Yi Mu, and Yongqing Li, editors. Cryptology and network security: 4th international conference, CANS 2005, Xiamen, China, December 14–16, 2005: proceedings, volume 3810 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30849-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I5534 2004.

URL http://www.springerlink.
com/content/978-3-540-
30849-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3810.

Dwork:2006:ACC

[Dwo06] Cynthia Dwork, editor.

Advances in Cryptology — CRYPTO 2006: 26th Annual International Cryptology Conference, Santa Barbara, California, USA, August 20–24, 2006. Proceedings, volume 4117 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37432-9 (print), 3-540-37433-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37433-6.

Dale:2005:NLP

[DWSK05] Robert Dale, Kam-Fai Wong, Jian Su, and Oi Yee Kwong, editors. Natural language processing, IJC-NLP 2005: second international joint conference, Jeju Island, Korea, October 11–13, 2005: proceedings, volume 3651 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29172-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 I372 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29172-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3651.

Deng:2009:AIC

[DWWL09] Hepu Deng, Lanzhou Wang, Fu Lee Wang, and Jingsheng Lei, editors. Artificial Intelligence and Computational Intelligence: International Conference, AICI 2009, Shanghai, China, November 7–8, 2009. Proceedings, volume 5855 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05252-5 (print), 3-642-05253-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
05253-8.

Dwyer:2001:MCS

[Dwy01] Matthew Dwyer, editor. Model checking software: 8th International SPIN Workshop, Toronto, Canada, May 19–20, 2001: Proceedings, volume 2057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42124-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 S65 2001; QA267.A1 L43 no.2057. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2057.htm;
http://www.springerlink.
com/content/978-3-540-
42124-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2057.

Ding:1991:STS

[DXS91] C. (Cunsheng) Ding, G. Xiao, and W. Shan.

The stability theory of stream ciphers, volume 561 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54973-0 (Berlin), 0-387-54973-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 187 pp. LCCN QA402.3 .D536 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0561.
htm;   http://www.springerlink.
com/content/978-3-540-
54973-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=561.

Dress:2007:COA

[DXZ07] Andreas Dress, Yinfeng Xu, and Binhai Zhu, editors. Combinatorial Optimization and Applications: First International Conference, COCOA 2007, Xi’an, China, August 14–16, 2007. Proceedings, volume 4616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73555-0 (print), 3-540-73556-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73556-4.

DePrisco:2006:SCN

[DY06a] Roberto De Prisco and Moti Yung, editors. Security and Cryptography for Networks: 5th International Conference, SCN 2006, Maiori, Italy, September 6–8, 2006. Proceedings, volume 4116 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38080-9 (print), 3-540-38081-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38081-8.

Deng:2006:INE

[DY06b] Xiaotie Deng and Yinyu Ye, editors. Internet and network economics: First International workshop, Wine 2005, Hong Kong, China, December 15–17, 2005. proceedings, volume 3828 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30900-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        30900-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3828.

DiNitto:2010:TSB

[DY10] Elisabetta Di Nitto and Ramin Yahyapour, editors. Towards a Service-Based Internet: Third European Conference, ServiceWave 2010, Ghent, Belgium, December 13–15, 2010. Proceedings, volume 6481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17693-3 (print), 3-642-17694-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17694-4.

Dyckhoff:1994:ELP

[Dyc94] Roy Dyckhoff, editor. Extensions of logic programming: 4th international workshop, ELP ’93, St Andrews, UK, March 29–April 1, 1993: proceedings, volume 798 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58025-5 (Berlin), 0-387-58025-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .E47 1993. DM72.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0798.
htm;   http://www.springerlink.
com/content/978-3-540-
58025-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=798.

Dyckhoff:2000:ARA

[Dyc00] Roy Dyckhoff, editor.

Automated reasoning with analytic tableaux and related methods: International Conference, TABLEAUX 2000, St Andrews, Scotland, UK, July 3–7, 2000: Proceedings, volume 1847 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67697-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1847. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1847.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67697-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1847.

Damiani:2009:AIB

[DYCD09] Ernesto Damiani, Kokou Yetongnon, Richard Chbeir, and Albert Dipanda, editors. Advanced Internet Based Systems and Applications: Second International Conference on Signal-Image Technology and Internet-Based Systems, SITIS 2006, Hammamet, Tunisia, December 17–21, 2006, Revised Selected Papers, volume 4879 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01349-X (print), 3-642-01350-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01350-8.

Du:1994:ACI

[DZ94] Dingzhu Du and Xiang-Sun Zhang, editors. Algorithms and computation: 5th International Symposium, ISAAC ’94, Beijing, P.R. China, August 25–27, 1994: proceedings, volume 834 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58325-4 (Berlin), 0-387-58325-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I73 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0834.htm;
 http://www.springerlink.
com/content/978-3-540-
58325-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=834.

DiBattista:2003:AEA

[DZ03] Giuseppe Di Battista and Uri Zwick, editors. Algorithms—ESA 2003: 11th annual European symposium, Budapest, Hungary, September 16– 19, 2003: Proceedings, volume 2832 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20064-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2832.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20064-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2832;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13632.

Dearle:2010:ODT

[DZ10a] Alan Dearle and Roberto V. Zicari, editors. Objects and Databases: Third International Conference, ICOODB 2010, Frankfurt/Main, Germany, September 28–30, 2010. Proceedings, volume 6348 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16091-3 (print), 3-642-16092-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16092-9.

Dong:2010:FMS

[DZ10b] Jin Song Dong and Huibiao Zhu, editors. Formal Methods and Software Engineering: 12th International Conference on Formal Engineering Methods, ICFEM 2010, Shanghai, China, November 17–19, 2010.

Proceedings, volume 6447 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16900-7 (print), 3-642-16901-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16901-4.

Dimitriadis:2006:GDI

[DZGS06] Yannis A. Dimitriadis, Ilze Zigurs, and Eduardo Gómez-Sánchez, editors. Groupware: Design, Implementation, and Use: 12th International Workshop, CRIWG 2006, Medina del Campo, Spain, September 17–21, 2006. Proceedings, volume 4154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39591-1 (print), 3-540-39595-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
39595-9.

Esparcia-Alcazar:2010:GPE

[EAES^{+}10] Anna Isabel Esparcia-Alcázar, Anikó Ekárt, Sara Silva, Stephen Dignum, and A. Şima Uyar, editors. Genetic Programming: 13th European Conference, EuroGP 2010, Istanbul, Turkey, April 7–9, 2010. Proceedings, volume 6021 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12147-0 (print), 3-642-12148-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12148-7.

Eckstein:2004:EPA

[EB04] Jutta Eckstein and Hubert Baumeister, editors. Extreme Programming and Agile Processes in Software Engineering: 5th International Conference, XP 2004, Garmisch-Partenkirchen, Germany, June 6–10, 2004. Proceedings, volume 3092 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22137-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 X62 2004.

URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t3092.
htm;   http://www.springerlink.
com/content/978-3-540-
22137-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3092;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98150.

Esposito:2008:VNF

[EBAH08] Anna Esposito, Nikolaos G. Bourbakis, Nikolaos Avouris, and Ioannis Hatzilygeroudis, editors. Verbal and Nonverbal Features of Human-Human and Human-Machine Interaction: COST Action 2102 International Conference, Patras, Greece, October 29–31, 2007. Revised Papers, volume 5042 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70871-5 (print), 3-540-70872-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70872-8.

Eiben:1998:PPS

[EBSS98] Agoston E. Eiben, Thomas Bäck, Marc Schoenauer, and Hans-Paul Schwefel, editors. Parallel problem solving from nature-PPSN V: 5th International Conference, Amsterdam, The Netherlands, September 27– 30, 1998: Proceedings, volume 1498 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65078-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1498. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1498.htm;
 http://www.springerlink.
com/content/978-3-540-
65078-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1498.

Enokido:2007:NBI

[EBT07] Tomoya Enokido, Leonard Barolli, and Makoto Takizawa, editors. Network-Based Information Systems: First International Conference, NBiS 2007, Regensburg, Germany, September 3–7, 2007. Proceedings, volume 4658 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74572-6 (print), 3-540-74573-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74573-0.

Eder:2011:ADI

[EBT11] Johann Eder, Maria Bielikova, and A. Min Tjoa, editors. Advances in Databases and Information Systems: 15th International Conference, ADBIS 2011, Vienna, Austria, September 20– 23, 2011. Proceedings, volume 6909 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23736-3 (print), 3-642-23737-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23737-9.

ElYacoubi:2006:CAI

[ECB06] Samira El Yacoubi, Bastien Chopard, and Stefania Bandini, editors. Cellular Automata: 7th International Conference on Cellular Automata, for Research and Industry, ACRI 2006, Perpignan, France, September 20–23, 2006. Proceedings, volume 4173 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-40929-7 (print), 3-540-40932-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
40932-8.

Esposito:2010:DMI

[ECV^{+}10] Anna Esposito, Nick Campbell, Carl Vogel, Amir Hussain, and Anton Nijholt, editors. Development of Multimodal Interfaces: Active Listening and Synchrony: Second COST 2102 International Training School, Dublin, Ireland, March 23–27, 2009, Revised Selected Papers, volume 5967 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12396-1 (print), 3-642-12397-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12397-9.

Eder:2006:BPM

[ED06a] Johann Eder and Schahram Dustdar, editors. Business Process Management Workshops: BPM 2006 International Workshops, BPD, BPI, ENEI, GPWW, DPM, semantics4ws, Vienna, Austria, September 4–7, 2006. Proceedings, volume 4103 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38444-8 (print), 3-540-38445-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38445-8.

Euzenat:2006:AIM

[ED06b] Jérôme Euzenat and John Domingue, editors. Artificial Intelligence: Methodology, Systems, and Applications: 12th International Conference, AIMSA 2006, Varna, Bulgaria, September 12– 15, 2006. Proceedings, volume 4183 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-40930-0 (print), 3-540-40931-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        40931-1.

Ehrig:2004:ISS

[EDD04] Hartmut Ehrig, Werner Damm, and Jörg Desel, editors. Integration of Software Specification Techniques for Applications in Engineering: priority program SoftSpez of the German Research Foundation (DFG), final report, volume 3147 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23135-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3147.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23135-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3147;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100778.

Ebert:2011:HAV

[EDGP11] Achim Ebert, Alan Dix, Nahum D. Gershon, and Margit Pohl, editors. Human Aspects of Visualization: Second IFIP WG 13.7 Workshop on Human-Computer Interaction and Visualization, HCIV (INTER-ACT) 2009, Uppsala, Sweden, August 24, 2009, Revised Selected Papers, volume 6431 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19640-3 (print), 3-642-19641-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19641-6.

Eigenmann:2008:ONE

[EdS08] Rudolf Eigenmann and Bronis R. de Supinski, editors. OpenMP in a New Era of Parallelism: 4th International Workshop, IWOMP 2008 West Lafayette, IN, USA, May 12–14, 2008 Proceedings, volume 5004 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79560-X (print), 3-540-79561-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79561-2.

Ehrig:1987:TPIa

[EEK^{+}87] Hartmut Ehrig, Ehrig, Robert Kowalski, Giorgio Levi, and Ugo Montanari, editors. TAPSOFT ’87: proceedings of the International Joint Conference on Theory and Practice of Software Development, Pisa, Italy, March 23–27, 1987. Volume 1. Advanced Seminar on Foundations of Innovative Software Development I and Colloquium on Trees in Algebra and Programming (CAAP ’87), volume 249 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17660-8, 3-540-17660-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I58 1987 v. 1–2 (c1987). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0249.htm;
 http://www.springerlink.
com/content/978-0-387-
17660-4;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=249.

Ehrig:2000:TAG

[EEKR00] Hartmut Ehrig, Gregor Engels, Hans-Jörg Kreowski, and Grzegorz Rozenberg, editors. Theory and application of graph transformations: 6th International Workshop, TAGT’98, Paderborn, Germany, November 16–20, 1998: Selected Papers, volume 1764 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67203-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1764. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1764.htm;
http://www.springerlink.
com/content/978-3-540-
67203-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1764.

Ehrig:2012:GTI

[EEKR12] Hartmut Ehrig, Gregor Engels, Hans-Jörg Kreowski, and Grzegorz Rozenberg, editors. Graph Transformations: 6th International Conference, ICGT 2012, Bremen, Germany, September 24–29, 2012. Proceedings, volume 7562 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33653-1 (print), 3-642-33654-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33654-6.

Eklund:1996:CSK

[EEM96] Peter W. Eklund, Gerard Ellis, and Graham Mann, editors. Conceptual structures: knowledge representation as interlingua: Fourth International Conference on Conceptual Structures, ICCS ’96, Sydney, Australia, August 19–22, 1996: proceedings, volume 1115 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61534-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2 .I58 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1115.

htm;   http://www.springerlink.
com/content/978-3-540-
61534-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1115.

Contents: Processes and participants / John F. Sowa — Conceptual structures of multicontexts / Rudolf Wille — Graph drawing methods / Peter Eades — Big posets of participatings and thematic roles / Fritz Lehmann

— Linguistic geometry tools generate optimal solutions / Boris Stilman — Natural language text processing and the maximal join operator / Heike Petermann — DB-MAT: knowledge acquisition, processing and NL generation using conceptual graphs / Galia Angelova and Kalina Bontcheva — Conceptual graph aboutness / Theo Huibers, Iadh Ounis and Jean-Pierre Chevallet — Conceptual structures and structured documents / Philippe Martin and Laurence Alpay — Conceptual graphs isomorphism: algorithm and use / Olivier Guinaldo — Editing conceptual graphs / Jens-Uwe Moller and Detlev Wiese — Computational situation theory in the conceptual graph language / Bikash Ghosh and Vila Wuwongse — Existential graphs and tense logic / Peter Ohrstrom — Microanalysis: acquiring database semantics in conceptual graphs / Harry S. Delugach and Thomas H. Hinke — Propagating truth and detecting contradiction in conceptual graph databases / John Esch and Robert Levinson — Sound and complete forward and backward chainings of graph rules / Eric Salvat and Marie-Laure Mugnier — Towards fuzzy conceptual graph programs / Vilas Wuwongse and Cao Hoang Tru — Sparse term encoding for dynamic taxonomies / Andrew Fall — Lattice embedding / Janos J. Sarbo — Local scaling in conceptual data systems / Gerd Stumme.

Esposito:2011:TAA

[EEM^{+}11] Anna Esposito, Antonietta M. Esposito, Raffaele Martone, Vincent C. Müller, and Gaetano Scarpetta, editors. Toward Autonomous, Adaptive, and Context-Aware Multimodal Interfaces. Theoretical and Practical Issues: Third COST 2102 International Training School, Caserta, Italy, March 15–19, 2010, Revised Selected Papers, volume 6456 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18183-X (print), 3-642-18184-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18184-9.

Erdi:2004:CNC

[ÉEMS04] Péter Érdi, Anna Esposito, Maria Marinaro, and Silvia Scarpetta, editors. Computational neuroscience: cortical dynamics: 8th International Summer School on Neural Nets, Erice, Italy, October 31–November 6, 2003, revised lectures, volume 3146 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22566-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP383 .I56 2003.

URL http://www.springerlink.
com/content/978-3-540-
22566-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3146;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99370.

Egly:2002:ARA

[EF02] Uwe Egly and Christian G. Fermüller, editors. Automated Reasoning with Analytic Tableaux and Related Methods: International Conference, TABLEAUX 2002. Copenhagen, Denmark, July 30–August 1, 2002. Proceedings, volume 2381 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43929-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T33 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2381.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43929-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2381.

Esik:2003:DLT

[EF03] Zoltan Esik and Zoltan Fulop, editors. Developments in language theory: 7th International Conference, DLT 2003, Szeged, Hungary, July 7–11, 2003: Proceedings, volume 2710 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40434-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .C63 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2710.htm;
 http://www.springerlink.
com/content/978-3-540-
40434-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2710.

Epstein:2012:AEA

[EF12] Leah Epstein and Paolo Ferragina, editors. Algorithms — ESA 2012: 20th Annual European Symposium, Ljubljana, Slovenia, September 10– 12, 2012. Proceedings, volume 7501 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33089-4 (soft cover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33089-6;   http://www.
                                        springerlink.com/content/
                                        978-3-642-33089-6/.

Ehrgott:2009:EMC

[EFG^{+}09] Matthias Ehrgott, Carlos M. Fonseca, Xavier Gandibleux, Jin-Kao Hao, and Marc Sevaux, editors. Evolutionary Multi-Criterion Optimization: 5th International Conference, EMO 2009, Nantes, France, April 7–10, 2009. Proceedings, volume 5467 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01019-9 (print), 3-642-01020-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01020-0.

Egenhofer:2004:GIS

[EFM04] Max J. Egenhofer, Christian Freksa, and Harvey J. Miller, editors. Geographic information science: third international conference, GI-Science 2004, Adelphi, MD, USA, October 20– 23, 2004, volume 3234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23558-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
23558-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3234;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101397.

Eusgeld:2008:DMA

[EFR08] Irene Eusgeld, Felix C.

Freiling, and Ralf Reussner, editors. Dependability Metrics: Advanced Lectures, volume 4909 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68946-X (print), 3-540-68947-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68947-8.

Eiter:2001:LPN

[EFT01] Thomas Eiter, Wolfgang Faber, and Miroslaw Truszczynski, editors. Logic programming and nonmonotonic reasoning:

6th International Conference, LPNMR 2001, Vienna, Austria, September 17–19, 2001: Proceedings, volume 2173 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42593-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .L632 2001; QA267.A1 L43 no.2173. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2173.htm;
 http://www.springerlink.
com/content/978-3-540-
42593-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2173.

Enjalbert:1993:SAS

[EFW93] P. Enjalbert, A. Finkel, and K. W. Wagner, editors. STACS 93: 10th Annual Symposium on Theoretical Aspects of Computer Science, Würzburg, Germany, February 25–27, 1993: proceedings, volume 665 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56503-5 (U.S.).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.665. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0665.htm;
http://www.springerlink.
com/content/978-0-387-
56503-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=665.

Esposito:2007:VNC

[EFZKM07] Anna Esposito, Marcos Faundez-Zanuy, Eric Keller, and Maria Marinaro, editors. Verbal and Nonverbal Communication Behaviours: COST Action 2102 International Workshop, Vietri sul Mare, Italy, March 29–31, 2007, Revised Selected and Invited Papers, volume 4775 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76441-0 (print), 3-540-76442-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76442-7.

Embley:1997:CME

[EG97] David W. Embley and Robert C. Goldstein, editors. Conceptual modeling — ER’97: 16th International Conference on Conceptual Modeling, Los Angeles, California, USA, November 3–5, 1997: proceedings, volume 1331 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63699-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26I567 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1331.htm;
 http://www.springerlink.
com/content/978-3-540-
63699-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1331.

Eppstein:2010:GDI

[EG10] David Eppstein and Emden R. Gansner, editors. Graph Drawing: 17th International Symposium, GD 2009, Chicago, IL, USA, September 22–25, 2009. Revised Papers, volume 5849 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11804-6 (print), 3-642-11805-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11805-0.

Egenhofer:2011:SIT

[EGMW11] Max Egenhofer, Nicholas Giudice, Reinhard Moratz, and Michael Worboys, editors. Spatial Information Theory: 10th International Conference, COSIT 2011, Belfast, ME, USA, September 12–16, 2011. Proceedings, volume 6899 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23195-0 (print), 3-642-23196-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23196-4.

Egges:2009:MGS

[EGO09] Arjan Egges, Roland Geraerts, and Mark Overmars, editors. Motion in Games: Second International Workshop, MIG 2009, Zeist, The Netherlands, November 21–24, 2009. Proceedings, volume 5884 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10346-4 (print), 3-642-10347-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10347-6.

Egenhofer:1995:ASD

[EH95] Max J. Egenhofer and John R. Herring, editors. Advances in spatial databases: 4th international symposium, SSD ’95, Portland, ME, USA, August 6–9, 1995: proceedings, volume 951 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60159-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.2 .S69 1995. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0951.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60159-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=951.

Eklund:2008:CSK

[EH08] Peter Eklund and Ollivier Haemmerlé, editors. Conceptual Structures: Knowledge Visualization and Reasoning: 16th International Conference on Conceptual Structures, ICCS 2008 Toulouse, France, July 7–11, 2008 Proceedings, volume 5113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70595-3 (print), 3-540-70596-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70596-3.

Even:2012:SIC

[EH12] Guy Even and Magnús M.

Halldórsson, editors. Structural Information and Communication Complexity: 19th International Colloquium, SIROCCO 2012, Reykjavik, Iceland, June 30–July 2, 2012, Revised Selected Papers, volume 7355 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31103-2 (print), 3-642-31104-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        31104-8.

Eleftherakis:2012:SEF

[EHH12] George Eleftherakis, Michael G. (Michael Gerard) Hinchey, and W. M. L. (William Michael Lloyd) Holcombe, editors. Software engineering and formal methods: 10th International Conference, SEFM 2012, Thessaloniki, Greece, October 1–5, 2012. Proceedings, volume 7504 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33825-9 (paperback), 3-642-33826-7 (ebook). LCCN QA76.758 .I58 2012. URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33825-0;   http://www.
                                        springerlink.com/content/
                                        978-3-642-33826-7.

Eder:2005:ADI

[EHKP05] Johann Eder, Hele-Mai Haav, Ahto Kalja, and Jaan Penjam, editors. Advances in Databases and Information Systems: 9th East European Conference, ADBIS 2005, Tallinn, Estonia, September 12–15, 2005, Proceedings, volume 3631 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28585-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28585-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3631.

Elomaa:2011:DSI

[EHM11] Tapio Elomaa, Jaakko Hollmén, and Heikki Mannila, editors. Discovery Science: 14th International Conference, DS 2011, Espoo, Finland, October 5–7, 2011. Proceedings, volume 6926 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24476-9 (print), 3-642-24477-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24477-3.

Esposito:2009:MSC

[EHMM09] Anna Esposito, Amir Hussain, Maria Marinaro, and Raffaele Martone, editors. Multimodal Signals: Cognitive and Algorithmic Issues: COST Action 2102 and euCognition International School Vietri sul Mare, Italy, April 21–26, 2008 Revised Selected and Invited Papers, volume 5398 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00524-1 (print), 3-642-00525-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00525-1.

Echtle:1994:DCE

[EHP94] Klaus Echtle, D. Hammer, and David Powell, editors. Dependable computing — EDCC-1: first European Dependable Computing Conference, Berlin, Germany, October 4–6, 1994: proceedings, volume 852 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58426-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F38 E33 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0852.
htm;   http://www.springerlink.
com/content/978-0-387-

58426-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=852.

Ehrig:1985:FMS

[Ehr85a] Hartmut Ehrig, editor.

Formal Methods and Software Development / Colloquium on Software Engineering (CSE). Proceedings of the International Joint Conference on Theory and Practice of Software Development (TAPSOFT), Berlin, March 25–29, 1985, volume 186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15199-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I57 1985 v. 1–2 (c1985). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0186.htm;
http://www.springerlink.
com/content/978-0-387-
15199-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=186.

Ehrig:1985:MFS

[Ehr85b] Hartmut Ehrig, editor.

Mathematical foundations of software development: Colloquium on Trees in Algebra and Programming (CAAP ’85). Proceedings of the International Joint Conference on Theory and Practice of Software Development (TAPSOFT), Berlin, March 25–29, 1985, volume 185 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15198-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I57 1985 v. 1–2 (c1985). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0185.htm;
 http://www.springerlink.
com/content/978-0-387-
15198-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=185.

Ehrig:1987:GGT

[Ehr87] Hartmut Ehrig, editor. Graph-grammars and their application to computer science: 3rd International Workshop, Warrenton, Virginia, USA, December 2–6, 1986: [proceedings], volume 291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18771-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .G681 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0291.htm;
 http://www.springerlink.
com/content/978-0-387-
18771-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=291.

Ehrig:1989:CMC

[Ehr89] Hartmut Ehrig, editor. Categorical methods in computer science: with aspects from topology, volume 393 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51722-7 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 C36 1989. DM52.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0393.
htm;   http://www.springerlink.
com/content/978-0-387-
51722-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=393.

Papers presented at the International Workshop on Categorical Methods in Computer Science with Aspects from Topology, held 1988 in Berlin.

Ehrig:1991:RTD

[Ehr91] Hartmut Ehrig, editor. Recent trends in data type specification: 7th Workshop on Specification of Abstract Data Types, Wusterhausen/ Dosse, Germany, April 17–20, 1990: proceedings, volume 534 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54496-8 (Berlin), 0-387-54496-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A23 W67 1990. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0534.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54496-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=534.

Ehrenberger:1998:CSR

[Ehr98] Wolfgang D. Ehrenberger, editor. Computer safety, reliability and security: 17th International Conference, SAFECOMP ’98, Heidelberg, Germany, October 5–7, 1998: Proceedings, volume 1516 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65110-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN A76.76.R47I585 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1516.htm;
 http://www.springerlink.
com/content/978-3-540-
65110-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1516.

Ehrig:2004:GTS

[Ehr04] Hartmut Ehrig, editor.

Graph transformations: second international conference, ICGT 2004, Rome, Italy, September 28–October 2, 2004: proceedings, volume 3256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23207-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .I24 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23207-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3256;
 http://www.springerlink.

                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100934.

Ehrig:2008:GTI

[EHRT08] Hartmut Ehrig, Reiko Heckel, Grzegorz Rozenberg, and Gabriele Taentzer, editors. Graph Transformations: 4th International Conference, ICGT 2008, Leicester, United Kingdom, September 7– 13, 2008. Proceedings, volume 5214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87404-6 (print), 3-540-87405-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87405-8.

Eriksson:1992:ELP

[EHSH92] L.-H. Eriksson, Lars Hallnas, and Peter Joseph Schroeder-Heister, editors. Extensions of logic programming: second international workshop, Stockholm, Sweden, January 27–29, 1991, proceedings, volume 596 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55498-X (Berlin), 0-387-55498-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .E98 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0596.htm;
 http://www.springerlink.
com/content/978-3-540-
55498-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=596.

Eskin:2006:SBR

[EIRW06] Eleazar Eskin, Trey Ideker, Ben Raphael, and Christopher Workman, editors. Systems Biology and Regulatory Genomics: Joint Annual RECOMB 2005 Satellite Workshops on Systems Biology and on Regulatory Genomics, San Diego, CA, USA; December 2–4, 2005, Revised Selected Papers, volume 4023 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48293-8 (print), 3-540-48540-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        48540-7.

Ebert:2010:MLB

[EJM10] Christian Ebert, Gerhard Jäger, and Jens Michaelis, editors. The Mathematics of Language: 10th and 11th Biennial Conference, MOL 10, Los Angeles, CA, USA, July 28–30, 2007, and MOL 11, Bielefeld, Germany, August 20–21, 2009, Revised Selected Papers, volume 6149 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14321-0 (print), 3-642-14322-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14322-9.

Ehrig:2001:UPN

[EJPR01] Harmut Ehrig, G. Juhás, J. Padberg, and G. Rozenberg, editors. Unifying Petri Nets: Advances in Petri Nets, volume 2128 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-43067-9. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 484 pp. LCCN QA267 .U54 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2128.htm;
http://www.springerlink.
com/content/978-3-540-
43067-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2128.

Etzion:1998:TDR

[EJS98] Opher Etzion, Sushil.

Jajodia, and Suryanarayana Sripada, editors. Temporal databases: research and practice, volume 1399 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64519-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1399.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1399.
htm;   http://www.springerlink.
com/content/978-3-540-
64519-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1399.

Ershov:1977:MAL

[EK77] A. P. Ershov and C. H. A.

Koster, editors. Methods of algorithmic language implementation, volume 47 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CODEN LNCSD9. ISBN 0-387-08065-1 (New York), 3-540-08065-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7.V7813 1975. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0047.htm;
 http://www.springerlink.
com/content/978-0-387-
08065-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=47.

English translation from the Russian of most of the proceedings (Trudy) of the symposium.

Ershov:1981:AMM

[EK81a] A. P. Ershov and Donald E. Knuth, editors. Algorithms in modern mathematics and computer science: proceedings, Urgench, Uzbek SSR, September 16–22, 1979, volume 122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-11157-3 (New York), 3-540-11157-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA9.58 .A43 1981. URL https:

//link.springer.com/
book/10.1007/3-540-

11157-3. The symposium was organized by the Academy of Sciences of the Uzbek S.S.R.

Even:1981:ALP

[EK81b] Shimon Even and O. Kariv, editors. Automata, languages, and programming: eighth colloquium, Acre (Akko), Israel, July 13–17, 1981, volume 115 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10843-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.115. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0115.htm;
http://www.springerlink.
com/content/978-0-387-
10843-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=115.

Sponsored by European Association for Theoretical Computer Science.

Erlebach:2007:AOA

[EK07] Thomas Erlebach and Christos Kaklamanis, editors. Approximation and Online Algorithms: 4th International Workshop, WAOA 2006, Zurich, Switzerland, September 14–15, 2006. Revised Papers, volume 4368 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69513-3 (print), 3-540-69514-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69514-1.

Edwards:2009:FFR

[EK09] Stephen H. Edwards and Gregory Kulczycki, editors. Formal Foundations of Reuse and Domain Engineering: 11th International Conference on Software Reuse, ICSR 2009, Falls Church, VA, USA, September 27– 30, 2009. Proceedings, volume 5791 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04210-4 (print), 3-642-04211-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04211-9.

Eliassen:2010:DAI

[EK10] Frank Eliassen and Rüdiger Kapitza, editors. Distributed Applications and Interoperable Systems: 10th IFIP WG 6.1 International Conference, DAIS 2010, Amsterdam, The Netherlands, June 7–9, 2010. Proceedings, volume 6115 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13644-3 (print), 3-642-13645-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13645-0.

Eiter:2012:RWS

[EK12] Thomas Eiter and Thomas Krennwallner, editors. Reasoning Web. Semantic Technologies for Advanced Query Answering: 8th International Summer School 2012, Vienna, Austria, September 3–8, 2012. Proceedings, volume 7487 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33157-2 (print), 3-642-33158-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33158-9.

Efthimiou:2012:GSL

[EKF12] Eleni Efthimiou, Georgios Kouroupetroglou, and Stavroula-Evita Fotinea, editors. Gesture and Sign Language in Human-Computer Interaction and Embodied Communication: 9th International Gesture Workshop, GW 2011, Athens, Greece, May 25–27, 2011, Revised Selected Papers, volume 7206 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34181-0 (print), 3-642-34182-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34182-3.

Eklundh:1994:CVEa

[Ekl94a] Jan-Olof Eklundh, editor. Computer vision, ECCV ’94: Third European Conference on Computer Vision, Stockholm, Sweden, May 2– 6, 1994: Proceedings. Part I, volume 800 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57956-7 (New York), 3-540-57956-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E97 1994 v.1–2 (c1994). DM114.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0800.htm;
 http://www.springerlink.
com/content/978-0-387-
57956-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=800.

Eklundh:1994:CVEb

[Ekl94b] Jan-Olof Eklundh, editor. Computer vision, ECCV ’94: Third European Conference on Computer Vision, Stockholm, Sweden, May 2– 6, 1994: Proceedings. Part II, volume 801 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57957-5 (New York), 3-540-57957-5 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E97 1994 v.1–2 (c1994). DM86.00. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0801.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57957-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=801.

Eklund:2004:CLS

[Ekl04] Peter Eklund, editor. Concept Lattices: Second International Conference on Formal Concept Analysis, ICFCA 2004, Sydney, Australia, February 23–26, 2004: Proceedings, volume 2961 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21043-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA171.5 .I24 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2961.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21043-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2961;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95548.

Eymann:2005:MST

[EKL^{+}05] Torsten Eymann, Franziska Klügl, Winfried Lamersdorf, Matthias Klusch, and Michael N. Huhns, editors. Multiagent system technologies: third German conference, MATES 2005, Koblenz, Germany, September 11– 13, 2005: proceedings, volume 3550 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28740-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28740-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3550.

Ehrig:1987:TPIb

[EKLM87] Hartmut Ehrig, Robert A.

Kowalski, Giorgio Levi, and Ugo Montanari, editors. TAPSOFT ’87: proceedings of the International Joint Conference on Theory and Practice of Software Development, Pisa, Italy, March 23–27, 1987. Volume 2. Advanced Seminar on Foundations of Innovative Software Development II and Colloqium on Functional and Logic Programming and Specifications (CFLP), volume 250 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17611-X, 3-540-17611-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN In Process. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0250.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        17611-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=250.

Etzion:2006:NGI

[EKM06] Opher Etzion, Tsvi Kuflik, and Amihai Motro, editors. Next Generation Information Technologies and Systems: 6th International Conference, NGITS 2006, Kibbutz Shefayim, Israel, July 4–6, 2006. Proceedings, volume 4032 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35472-7 (print), 3-540-35473-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
35473-4.

Enser:2004:IVR

[EKO^{+}04] Peter (G. B.) Enser, Yiannis Kompatsiaris, Noel E. O’Connor, Alan F. Smeaton, and Arnold W. M. Smeulders, editors. Image and Video Retrieval: Third International Conference, CIVR 2004, Dublin, Ireland, July 21–23, 2004. Proceedings, volume 3115 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22539-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .C49 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3115.htm;
http://www.springerlink.
com/content/978-3-540-
22539-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3115;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98923.

Egges:2008:MGF

[EKO08] Arjan Egges, Arno Kamphuis, and Mark Overmars, editors. Motion in Games: First International Workshop, MIG 2008, Utrecht, The Netherlands, June 14– 17, 2008. Revised Papers, volume 5277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89219-2 (print), 3-540-89220-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89220-5.

Eleftherakis:2007:MCI

[EKP^{+}07] George Eleftherakis, Petros Kefalas, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa, editors. Membrane Computing: 8th International Workshop, WMC 2007 Thessaloniki, Greece, June 25–28, 2007 Revised Selected and Invited Papers, volume 4860 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77311-8 (print), 3-540-77312-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77312-2.

Ehrig:1991:GGT

[EKR91] Hartmut Ehrig, H.-J.

Kreowski, and Grzegorz Rozenberg, editors. Graph grammars and their application to computer science: 4th International Workshop, Bremen, Germany, March 5–9, 1990: proceedings, volume 532 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54478-X (Berlin), 0-387-54478-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .G68 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0532.htm;
 http://www.springerlink.
com/content/978-3-540-
54478-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=532.

ElGayar:2010:MCS

[EKR10] Neamat El Gayar, Josef Kittler, and Fabio Roli, editors. Multiple Classifier Systems: 9th International Workshop, MCS 2010, Cairo, Egypt, April 7–9, 2010. Proceedings, volume 5997 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12126-8 (print), 3-642-12127-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12127-2.

Evans:2000:FUU

[EKS00] Andy Evans, Stuart (J. H.) Kent, and Bran Selic, editors. ^{≪}UML^{≫}2000— The Unified Modeling Language. Advancing the Standard: Third International Conference, York, UK, October 2–6, 2000. Proceedings, volume 1939 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41133-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1939. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1939.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41133-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1939.

Echizen:2010:AIC

[EKS10] Isao Echizen, Noboru Kunihiro, and Ryoichi Sasaki, editors. Advances in Information and Computer Security: 5th International Workshop on Security, IWSEC 2010, Kobe, Japan, November 22–24, 2010. Proceedings, volume 6434 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16824-8 (print), 3-642-16825-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16825-3.

Elmasri:1994:ERA

[EKT94] Ramez Elmasri, Vram Kouramajian, and B. Thalheim, editors. Entityrelationship approach, ER ’93: 12th International Conference on the Entity-Relationship Approach, Arlington, Texas, USA, December 15–17, 1993: proceedings, volume 823 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58217-7 (New York), 3-540-58217-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.9.D3 I5584 1993. DM96.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0823.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        58217-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=823.

Eriksson:2002:FFM

[EL02a] L.-H. Eriksson and Peter Alexander Lindsay, editors. FME 2002: Formal Methods—Getting IT Right: International Symposium of Formal Methods Europe, Copenhagen, Denmark, July 22–24, 2002. Proceedings, volume 2391 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43928-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I593 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2391.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43928-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2391.

Esparza:2002:ATP

[EL02b] Javier Esparza and Charles Lakos, editors. Application and Theory of Petri Nets 2002: 23rd International Conference, ICATPN 2002, Adelaide, Australia, June 24–30, 2002. Proceedings, volume 2360 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43787-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A66 2002 Bar. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2360.htm;
 http://www.springerlink.
com/content/978-3-540-
43787-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2360.

Eiter:2004:DTI

[EL04] Thomas Eiter and Leonid Libkin, editors. Database Theory: ICDT 2005: 10th international conference, Edinburgh, UK, January 5–7, 2005: Proceedings, volume 3363 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24288-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I557 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        24288-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3363;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104421.

Edelkamp:2007:MCA

[EL07] Stefan Edelkamp and Alessio Lomuscio, editors. Model Checking and Artificial Intelligence: 4th Workshop, MoChArt IV, Riva del Garda, Italy, August 29, 2006, Revised Selected and Invited Papers, volume 4428 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74127-5 (print), 3-540-74128-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74128-2.

Ellis:1995:CSA

[Ell95] Gerard Ellis, editor. Conceptual structures: applications, implementation, and theory: Third International Conference on Conceptual Structures, ICCS ’95, Santa Cruz, CA, USA, August 14–18, 1995: proceedings, volume 954 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60161-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2 .I58 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0954.
htm;   http://www.springerlink.
com/content/978-3-540-
60161-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=954.

Erdem:2012:CRE

[ELLP12] Esra Erdem, Joohyung Lee, Yuliya Lierler, and David Pearce, editors. Correct Reasoning: Essays on Logic-Based AI in Honour of Vladimir Lifschitz, volume 7265 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30742-6 (print), 3-642-30743-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30743-0.

Eigenmann:2005:LCH

[ELM05] Rudolf Eigenmann, Zhiyuan Li, and Samuel P. Midkiff, editors. Languages and compilers for high performance computing: 17th international workshop, LCPC 2004, West Lafayette, IN, USA, September 22– 24, 2004: revised selected papers, volume 3602 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28009-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .W67 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28009-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3602.

Elmoataz:2010:ISP

[ELN^{+}10] Abderrahim Elmoataz, Olivier Lezoray, Fathallah Nouboud, Driss Mammass, and Jean Meunier, editors. Image and Signal Processing: 4th International Conference, ICISP 2010, Trois-Rivières, QC, Canada, June 30–July 2, 2010. Proceedings, volume 6134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13680-X (print), 3-642-13681-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13681-8.

Elmoataz:2008:ISP

[ELNM08] Abderrahim Elmoataz, Olivier Lezoray, Fathallah Nouboud, and Driss Mammass, editors. Image and Signal Processing: 3rd International Conference, ICISP 2008. Cherbourg-Octeville, France, July 1 — 3, 2008. Proceedings, volume 5099 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69904-X (print), 3-540-69905-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69905-7.

Erdem:2009:LPN

[ELS09] Esra Erdem, Fangzhen Lin, and Torsten Schaub, editors. Logic Programming and Nonmonotonic Reasoning: 10th International Conference, LP-NMR 2009, Potsdam, Germany, September 14– 18, 2009. Proceedings, volume 5753 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04237-6 (print), 3-642-04238-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04238-6.

Engels:2010:GTM

[ELS^{+}10] Gregor Engels, Claus Lewerentz, Wilhelm Schäfer, Andy Schürr, and Bernhard Westfechtel, editors. Graph Transformations and Model-Driven Engineering: Essays Dedicated to Manfred Nagl on the Occasion of his 65th Birthday, volume 5765 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17321-7 (print), 3-642-17322-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17322-6.

Eder:2012:HSV

[ELS12] Kerstin Eder, João Lourenço, and Onn Shehory, editors. Hardware and Software: Verification and Testing: 7th International Haifa Verification Conference, HVC 2011, Haifa, Israel, December 6–8, 2011, Revised Selected Papers, volume 7261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34187-X (print), 3-642-34188-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34188-5.

Egenhofer:2002:GIS

[EM02] Max J. Egenhofer and David M. Mark, editors. Geographic Information Science: Second International Conference, GI-Science 2002, Boulder, CO, USA, September 25–28, 2002. Proceedings, volume 2478 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44253-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.212 .I5846 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2478.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44253-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2478.

Eder:2003:AIS

[EM03] J. Eder and M. Missikoff, editors. Advanced information systems engineering: 15th International Conference, CAiSE 2003, Klagenfurt/Velden, Austria, June 16–18, 2003: Proceedings, volume 2681 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40442-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .N67 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2681.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40442-2;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2681.

Eisinger:2005:RWF

[EM05] Norbert Eisinger and Jan Małuszynski, editors. Reasoning Web: First International Summer School 2005, Msida, Malta, July 25–29, 2005, Revised Lectures, volume 3564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27828-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815 .S89 2005.

URL http://www.springerlink.
com/content/978-3-540-
27828-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3564.

Eliassen:2006:DAI

[EM06] Frank Eliassen and Alberto Montresor, editors. Distributed Applications and Interoperable Systems: 6th IFIP WG 6.1 International Conference, DAIS 2006, Athens, Greece, June 14–15, 2006. Proceedings, volume 4025 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35126-4 (softcover).

LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        35126-9.

Egerstedt:2008:HSC

[EM08] Magnus Egerstedt and Bud Mishra, editors. Hybrid Systems: Computation and Control: 11th International Workshop, HSCC 2008, St. Louis, MO, USA, April 22– 24, 2008. Proceedings, volume 4981 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78928-6 (print), 3-540-78929-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78929-1.

Esparza:2010:TAC

[EM10] Javier Esparza and Rupak Majumdar, editors. Tools and Algorithms for the Construction and Analysis of Systems: 16th International Conference, TACAS 2010, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2010, Paphos, Cyprus, March 20–28, 2010. Proceedings, volume 6015 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12001-6 (print), 3-642-12002-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12002-2.

Embury:1998:ADB

[Emb98] Suzanne M. Embury, editor. Advances in databases: 16th British National Conference on Databases, BNCOD 16, Cardiff, Wales, UK, July 6–8, 1998: Proceedings, volume 1405 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64659-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1405. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1405.htm;
 http://www.springerlink.
com/content/978-3-540-
64659-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1405.

Elmoataz:2012:ISP

[EML^{+}12] Abderrahim Elmoataz, Driss Mammass, Olivier Lezoray, Fathallah Nouboud, and Driss Aboutajdine, editors. Image and Signal Processing: 5th International Conference, ICISP 2012, Agadir, Morocco, June 28–30, 2012. Proceedings, volume 7340 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31253-5 (print), 3-642-31254-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31254-0.

Elomaa:2010:AAE

[EMO10] Tapio Elomaa, Heikki Mannila, and Pekka Orponen, editors. Algorithms and Applications: Essays Dedicated to Esko Ukkonen on the Occasion of His 60th Birthday, volume 6060 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12475-5 (print), 3-642-12476-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12476-1.

Elomaa:2002:MLE

[EMT02a] Tapio Elomaa, Heikki Mannila, and Hannu Toivonen, editors. Machine Learning: ECML 2002: 13th European Conference on Machine Learning, Helsinki, Finland, August 19–23, 2002. Proceedings, volume 2430 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44036-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2430.htm;
http://www.springerlink.
com/content/978-3-540-
44036-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2430.

Elomaa:2002:PDM

[EMT02b] Tapio Elomaa, Heikki Mannila, and Hannu Toivonen, editors. Principles of Data Mining and Knowledge Discovery: 6th European Conference, PKDD 2002, Helsinki, Finland, August 19–23, 2002. Proceedings, volume 2431 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44037-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 P16 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2431.htm;
 http://www.springerlink.
com/content/978-3-540-
44037-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2431.

Enjalbert:1994:SAS

[EMW94] P. Enjalbert, Ernst Mayr, and K. W. Wagner, editors. STACS 94: 11th Annual Symposium on Theoretical Aspects of Computer Science, Caen, France, February 1994: proceedings, volume 775 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57785-8 (Berlin), 0-387-57785-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S956 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0775.htm;
http://www.springerlink.
com/content/978-3-540-
57785-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=775.

Ershov:1974:IST

[EN74] A. P. (Andreí Petrovich)

Ershov and Valerii Aleksandrovich Nepomniashchii, editors. International Symposium on Theoretical Programming [Proceedings held on August 7–11, 1972], volume 5 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 0-387-06720-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.5. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0005.htm;
http://www.springerlink.
com/content/978-0-387-
06720-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=5.

Emerson:2006:VMC

[EN06] E. Allen Emerson and Kedar S. Namjoshi, editors. Verification, Model Checking, and Abstract Interpretation: 7th International Conference, VMCAI 2006, Charleston, SC, USA, January 8–10, 2006. Proceedings, volume 3855 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31139-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31139-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3855.

Encarnacao:1980:CAD

[Enc80] Jose Luis Encarnacao, editor. Computer aided design: modelling, systems engineering, CADsystems: CREST Advanced Course, Darmstadt, September 1980, volume 89 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10242-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.89. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t0089.htm;
http://www.springerlink.
com/content/978-0-387-
10242-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=89.

Engelfriet:1974:SPS

[Eng74] Joost Engelfriet. Simple program schemes and formal languages, volume 20 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 0-387-06953-4. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 254 pp. LCCN QA267.3 .E531. DM25.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0020.htm;
http://www.springerlink.
com/content/978-0-387-
06953-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=20.

Engeler:1981:LPW

[Eng81] Erwin Engeler, editor.

Logic of programs: workshop, ETH Zürich, May– July 1979, volume 125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-11160-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L586 1981. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0125.htm;
 http://www.springerlink.
com/content/978-0-387-
11160-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=125.

Erlebach:2012:ASS

[ENO12] Thomas Erlebach, Sotiris Nikoletseas, and Pekka Orponen, editors. Algorithms for Sensor Systems: 7th International Symposium on Algorithms for Sensor Systems, Wireless Ad Hoc Networks and Autonomous Mobile Entities, ALGOSENSORS 2011, Saarbrücken, Germany, September 8–9, 2011, Revised Selected Papers, volume 7111 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28208-3 (print), 3-642-28209-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28209-6.

Eaglestone:2002:ADB

[ENP02] Barry Eaglestone, Siobhán North, and Alexandra Poulovassilis, editors. Advances in Databases: 19th British National Conference on Databases, BNCOD 19, Sheffield, UK, July 17– 19, 2002. Proceedings, volume 2405 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43905-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B75 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2405.htm;
http://www.springerlink.
com/content/978-3-540-
43905-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2405.

Ehrig:1983:GGT

[ENR83] Hartmut Ehrig, Manfred Nagl, and Grzegorz Rozenberg, editors. Graph-grammars and their application to computer science: 2nd International Workshop, Haus Ohrbeck, Germany, October 4– 8, 1982, volume 153 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-12310-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .G68 1983. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0153.htm;
 http://www.springerlink.
com/content/978-0-387-
12310-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=153.

Ehrig:1994:RTD

[EO94] Hartmut Ehrig and F. Orejas, editors. Recent trends in data type specification: 9th Workshop on Specification of Abstract Data Types, joint with the 4th COMPASS Workshop, Caldes de Malavella, Spain, October 26–30, 1992: Selected Papers, volume 785 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57867-6 (Berlin), 0-387-57867-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A23 W67 1992. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t0785.htm;
http://www.springerlink.
com/content/978-3-540-
57867-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=785.

Ebner:2007:GPE

[EOE^{+}07] Marc Ebner, Michael O’Neill, Anikó Ekárt, Leonardo Vanneschi, and Anna Isabel Esparcia-Alcázar, editors. Genetic Programming: 10th European Conference, EuroGP 2007, Valencia, Spain, April 11– 13, 2007. Proceedings, volume 4445 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71602-5 (print), 3-540-71605-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71605-1.

Obbink:2005:SPL

[eOP05] enk Obbink and Klaus Pohl, editors. Software product lines: 9th international conference, SPLC 2005, Rennes, France, September 26– 29, 2005, proceedings, volume 3714 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28936-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28936-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3714.

Embley:2006:CMI

[EOR06] David W. Embley, Antoni Olivé, and Sudha Ram, editors. Conceptual Modeling — ER 2006: 25th International Conference on Conceptual Modeling, Tucson, AZ, USA, November 6– 9, 2006. Proceedings, volume 4215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47224-X (print), 3-540-47227-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47227-8.

Engels:2007:MDE

[EOSW07] Gregor Engels, Bill Opdyke, Douglas C. Schmidt, and Frank Weil, editors. Model Driven Engineering Languages and Systems: 10th International Conference, MoDELS 2007, Nashville, USA, September 30 — October 5, 2007. Proceedings, volume 4735 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75208-0 (print), 3-540-75209-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75209-7.

Ellis:2003:MICa

[EP03a] Randy E. Ellis and Terry M. Peters, editors. Medical Image Computing and Computer-Assisted Intervention— MICCAI 2003: 6th International Conference, Montréal, Canada, November 15–18, 2003: Proceedings. Part 1, volume 2878 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20462-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R858.A2 I575 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2878.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20462-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2878;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b93810.

Ellis:2003:MICb

[EP03b] Randy E. Ellis and Terry M. Peters, editors. Medical Image Computing and Computer-Assisted Intervention— MICCAI 2003: 6th International Conference, Montréal, Canada, November 15–18, 2003: Proceedings. Part 2, volume 2879 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20464-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R858.A2 I575 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2879.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20464-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2879;

http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b93811.

Erlebach:2006:AOA

[EP06] Thomas Erlebach and Giuseppe Persinao, editors. Approximation and Online Algorithms: Third International Workshop, WAOA 2005, Palma de Mallorca, Spain, October 6–7, 2005, Revised Selected Papers, volume 3879 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32207-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32207-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3879.

Ersboll:2007:IAS

[EP07] Bjarne Kjær Ersbøll and Kim Steenstrup Pedersen, editors. Image Analysis: 15th Scandinavian Conference, SCIA 2007, Aalborg, Denmark, June 10–14, 2007, volume 4522 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73039-7 (print), 3-540-73040-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73040-8.

England:2010:TMD

[EPVW10] David England, Philippe Palanque, Jean Vanderdonckt, and Peter J. Wild, editors. Task Models and Diagrams for User Interface Design: 8th International Workshop, TAMODIA 2009, Brussels, Belgium, September 23–25, 2009, Revised Selected Papers, volume 5963 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11796-1 (print), 3-642-11797-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11797-8.

Etessami:2005:CAV

[ER05] Kousha Etessami and Sriram K. Rajamani, editors. Computer aided verification: 17th international conference, CAV 2005, Edinburgh, Scotland, UK, July 6– 10, 2005: proceedings, volume 3576 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27231-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E93 C38 2005. URL http:

//www.springerlink.
com/content/978-3-540-
27231-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3576.

Even:2012:DAA

[ER12] Guy Even and Dror Rawitz, editors. Design and Analysis of Algorithms: First Mediterranean Conference on Algorithms, MedAlg 2012, Kibbutz Ein Gedi, Israel, December 3–5, 2012. Proceedings, volume 7659 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34861-0 (print), 3-642-34862-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34862-4.

Elgammal:2007:HMU

[ERK07] Ahmed Elgammal, Bodo Rosenhahn, and Reinhard Klette, editors. Human Motion — Understanding, Modeling, Capture and Animation: Second Workshop, Human Motion 2007, Rio de Janeiro, Brazil, October 20, 2007. Proceedings, volume 4814 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75702-3 (print), 3-540-75703-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75703-0.

Erlebach:2006:CAA

[Erl06] Thomas Erlebach, editor. Combinatorial and Algorithmic Aspects of Networking: Third Workshop, CAAN 2006, Chester, UK, July 2, 2006. Revised Papers, volume 4235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48822-7 (print), 3-540-48824-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48824-8.

Esposito:2006:FIS

[ERMS06] Floriana Esposito, Zbigniew W. Raś, Donato Malerba, and Giovanni Semeraro, editors. Foundations of Intelligent Systems: 16th International Symposium, IS-MIS 2006, Bari, Italy, September 27–29, 2006. Proceedings, volume 4203 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45764-X (print), 3-540-45766-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45766-4.

Ernst:2007:EOO

[Ern07] Erik Ernst, editor. ECOOP 2007 — Object-Oriented Programming: 21st European Conference, Berlin, Germany, July 30 — August 3, 2007. Proceedings, volume 4609 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73588-7 (print), 3-540-73589-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73589-2.

Ehrig:2010:GTI

[ERRS10] Hartmut Ehrig, Arend Rensink, Grzegorz Rozenberg, and Andy Schürr, editors. Graph Transformations: 5th International Conference, ICGT 2010, Enschede, The Netherlands, September 27–October 2, 2010. Proceedings, volume 6372 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15927-3 (print), 3-642-15928-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15928-2.

Ehrig:2003:PNT

[ERRW03] Hartmut Ehrig, Wolfgang Reisig, Grzegorz Rozenberg, and Herbert Weber, editors. Petri Net Technology for Communication-Based Systems: Advances in Petri Nets, volume 2472 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20538-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .P49 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2472.htm;
 http://www.springerlink.
com/content/978-3-540-
20538-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2472;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b83526.

Eder:1999:ADI

[ERW99] J. (Johann) Eder, Ivan Rozman, and Tatjana Welzer, editors. Advances in databases and information systems: Third East European Conference, ADBIS’99, Maribor, Slovenia, September 13–16, 1999: Proceedings, volume 1691 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66485-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 A3396 1999. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t1691.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66485-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1691.

Etiemble:1992:PPA

[ES92] D. Etiemble and J.-C. Syre, editors. PARLE ’92, parallel architectures and languages Europe: 4th International PARLE Conference, Paris, France, June 15–18, 1992, proceedings, volume 605 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55599-4 (Berlin), 0-387-55599-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I564 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0605.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55599-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=605.

Emerson:2000:CAV

[ES00a] E. Allen Emerson and Aravinda Prasad Sistla, editors. Computer aided verification: 12th International Conference, CAV 2000, Chicago, IL, USA, July 15–19, 2000: Proceedings, volume 1855 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67770-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1855.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1855.
htm;   http://www.springerlink.
com/content/978-3-540-
67770-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1855.

Etzion:2000:CIS

[ES00b] Opher Etzion and Peter Scheuermann, editors. Cooperative information systems: 7th International Conference, CoopIS 2000, Eilat, Israel, September 6– 8, 2000: Proceedings, volume 1901 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41021-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I158 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1901.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41021-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1901.

Eiter:2002:FIK

[ES02] Thomas Eiter and Klaus-Dieter Schewe, editors. Foundations of Information and Knowledge Systems: Second International Symposium, FoIKS 2002, Salzau Castle, Germany, February 20–23, 2002: Proceedings, volume 2284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43220-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .F65 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2284.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43220-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2284.

Endler:2003:MAI

[ES03] Markus Endler and Douglas C. Schmidt, editors. Middleware 2003: ACM/ IFIP/USENIX International middleware Conference, Rio de Janeiro, Brazil, June 16–20, 2003: Proceedings, volume 2672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40317-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2672.htm;
 http://www.springerlink.
com/content/978-3-540-
40317-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2672.

Eisenbrand:2010:IPC

[ES10] Friedrich Eisenbrand and F. Bruce Shepherd, editors. Integer Programming and Combinatorial Optimization: 14th International Conference, IPCO 2010, Lausanne, Switzerland, June 9– 11, 2010. Proceedings, volume 6080 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13035-6 (print), 3-642-13036-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13036-6.

Escobar:2010:FCL

[Esc10] Santiago Escobar, editor. Functional and Constraint Logic Programming: 18th International Workshop, WFLP 2009, Brasilia, Brazil, June 28, 2009, Revised Selected Papers, volume 5979 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11998-0 (print), 3-642-11999-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11999-6.

Esik:1993:FCT

[Esi93] Zoltan Esik, editor. Fundamentals of computation theory: 9th international conference, FCT ’93, Szeged, Hungary, August 23–27, 1993: proceedings, volume 710 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57163-9 (Berlin: softcover), 0-387-57163-9 (New York: softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I5455 1993. DM86.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0710.htm;
http://www.springerlink.
com/content/978-3-540-
57163-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=710.

Esik:2006:CSL

[Ési06] Zoltán Ésik, editor.

Computer Science Logic: 20th International Workshop, CSL 2006, 15th Annual Conference of the EACSL, Szeged, Hungary, September 25– 29, 2006. Proceedings, volume 4207 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45458-6 (print), 3-540-45459-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45459-5.

Esposito:2001:AAA

[Esp01] Floriana Esposito, editor. AI*IA 2001: advances in artificial intelligence: 7th Congress of the Italian Association for Artificial Intelligence, Bari, Italy, September 25–28, 2001: Proceedings, volume 2175 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42601-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A88 2001; QA267.A1 L43 no.2175. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2175.htm;
 http://www.springerlink.
com/content/978-3-540-
42601-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2175.

Estublier:1995:SCM

[Est95] Jacky Estublier. Software configuration management: ICSE SCM-4 and SCM-5 workshops: Selected Papers, volume 1005 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60578-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 309 pp. LCCN QA76.76.C69 S64 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1005.
htm;   http://www.springerlink.
com/content/978-3-540-
60578-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1005.

Estesiano:1998:FAS

[Est98] E. (Egidio) Estesiano, editor. Fundamental approaches to software engineering: First International Conference, FASE’98 held as part of the joint European Conferences on Theory and Practice of Software, ETAPS’98, Lisbon, Portugal, March 28– April 4, 1998: Proceedings, volume 1382 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64303-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758.F37 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1382.htm;
 http://www.springerlink.
com/content/978-3-540-
64303-6;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1382.

Estublier:1999:SCM

[Est99] Jacky Estublier, editor. System configuration management: 9th International Symposium, SCM-9, Toulouse, France, September 5– 7, 1999: Proceedings, volume 1675 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66484-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1675; QA76.76.C69 S37 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1675.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66484-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1675.

Eades:2001:ACI

[ET01a] Peter Eades and Tadao Takaoka, editors. Algorithms and Computation: 12th International Symposium, ISAAC 2001, Christchurch, New Zealand, December 19–21, 2001. Proceedings, volume 2223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42985-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I73 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2223.
htm;   http://www.springerlink.
com/content/978-3-540-
42985-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2223.

Emmerich:2001:EDO

[ET01b] Wolfgang Emmerich and Stefan Tai, editors. Engineering distributed objects: Second International Workshop, EDO 2000, Davis, CA, USA, November 2–3, 2000: Revised Papers, volume 1999 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41792-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1999; QA76.64 .E25 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1999.htm;
 http://www.springerlink.
com/content/978-3-540-
41792-7;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1999.

Etalle:2006:LPI

[ET06] Sandro Etalle and Mirosław Truszczyński, editors. Logic Programming: 22nd International Conference, ICLP 2006, Seattle, WA, USA, August 17–20, 2006. Proceedings, volume 4079 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36635-0 (print), 3-540-36636-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        36636-2.

Etalle:2005:LBP

[Eta05] Sandro Etalle, editor. Logic based program synthesis and transformation: 14th international symposium, LOPSTR 2004, Verona, Italy, August 26–28, 2004: revised selected papers, volume 3573 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26655-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I565 2004. URL http:

//www.springerlink.
com/content/978-3-540-
26655-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3573.

Escrig:2002:TAI

[ETG02] M. T. Escrig, F. Toledo, and E. Golobardes, editors. Topics in Artificial Intelligence: 5th Catalonian Conference on AI, CCIA 2002, Castel·lon, Spain, October 24–25, 2002. Proceedings, volume 2504 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00011-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .C37 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2504.htm;
http://www.springerlink.
com/content/978-3-540-
00011-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2504.

Eugster:2009:OOT

[Eug09] Patrick Eugster, editor.

Object-Oriented Technology. ECOOP 2008 Workshop Reader: ECOOP 2008 Workshops Paphos, Cyprus, July 7– 11, 2008 Final Reports, volume 5475 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02046-1 (print), 3-642-02047-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02047-6.

Ebeling:1996:PPS

[EV96] Werner Ebeling and Hans-Michael Voigt, editors. Parallel problem solving from nature— PPSN IV: International Conference on Evolutionary Computation, the 4th Conference on Parallel Problem Solving from Nature, Berlin, Germany, September 22– 26, 1996: proceedings, volume 1141 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61723-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I535 1996. URL

http://link.springer-
ny.com/link/service/

series/0558/tocs/t1141.
htm;   http://www.springerlink.
com/content/978-3-540-
61723-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1141.

Eigenmann:2001:OSM

[EV01] Rudolf Eigenmann and Michael J. Voss, editors. OpenMP shared memory parallel programming: International Workshop on OpenMP Applications and Tools, WOMPAT 2001, West Lafayette, IN, USA, July 30–31, 2001: Proceedings, volume 2104 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42346-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .I589 2001; QA267.A1 L43 no.2104. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2104.htm;
http://www.springerlink.
com/content/978-3-540-
42346-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2104.

Escolano:2007:GBR

[EV07] Francisco Escolano and Mario Vento, editors.

Graph-Based Representations in Pattern Recognition: 6th IAPR-TC-15 International Workshop, GbRPR 2007, Alicante, Spain, June 11–13, 2007. Proceedings, volume 4538 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72902-X (print), 3-540-72903-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72903-7.

Esposito:2009:CMA

[EV09] Anna Esposito and Robert Vích, editors. Cross-Modal Analysis of Speech, Gestures, Gaze and Facial Expressions: COST Action 2102 International Conference Prague, Czech Republic, October 15–18, 2008 Revised Selected and Invited Papers, volume 5641 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03319-9 (print), 3-642-03320-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
03320-9.

Everling:1975:ECS

[Eve75] Wolfgang Everling. Exercises in computer systems analysis, volume 35 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, corrected reprint of the first edition, 1975. CODEN LNCSD9. ISBN 0-387-07401-5. ISSN 0302-9743 (print), 1611-3349 (electronic). iv + 4 + 184 pp. LCCN T57.5 .E84 1975. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0035.htm;
http://www.springerlink.
com/content/978-0-387-
07401-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=35.

Esposito:2011:AVN

[EVV^{+}11] Anna Esposito, Alessandro Vinciarelli, Klára Vicsi, Catherine Pelachaud, and Anton Nijholt, editors. Analysis of Verbal and Nonverbal Communication and Enactment. The Processing Issues: COST 2102 International Conference, Budapest, Hungary, September 7–10, 2010, Revised Selected Papers, volume 6800 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25774-7 (print), 3-642-25775-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25775-9.

Endres:1991:SDE

[EW91] Albert Endres and Herbert Weber, editors. Software development environments and CASE technology: European symposium, Konigswinter, June 17–19, 1991 proceedings, volume 509 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54194-2 (Berlin), 0-387-54194-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 S637 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0509.htm;
 http://www.springerlink.
com/content/978-3-540-
54194-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=509.

Erdogmus:2003:CBS

[EW03] Hakan Erdogmus and Tao Weng, editors. COTSbased software systems: Second International Conference, ICCBSS 2003, Ottawa, Canada, February 10–12, 2003: Proceedings, volume 2580 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00562-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .I25 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2580.htm;
http://www.springerlink.
com/content/978-3-540-
00562-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2580.

Also available via the World Wide Web.

Emmerich:2004:CDS

[EW04] Wolfgang Emmerich and Alexander L. Wolf, editors. Component Deployment: Second International Working Conference, CD 2004, Edinburgh, UK, May 20– 21, 2004: Proceedings, volume 3083 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22059-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C66 W. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3083.
htm;   http://www.springerlink.
com/content/978-3-540-
22059-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3083;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98000.

Eskin:2005:RGR

[EW05] Eleazar Eskin and Chris Workman, editors. Regulatory genomics: RE-COMB 2004 international workshop, RRG 2004, San Diego, CA., USA, March 26–27, 2004: revised selected papers, volume 3318 of Lecture Notes in Bioinformatics and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24456-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH450 .R74 2004. URL

http://springerlink.
metapress.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3318;
http://www.springerlink.
com/content/978-3-540-
24456-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3318;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105490.

Erlingsson:2011:ESS

[EWZ11] Úlfar Erlingsson, Roel Wieringa, and Nicola Zannone, editors. Engineering Secure Software and Systems: Third International Symposium, ESSoS 2011, Madrid, Spain, February 9–10, 2011. Proceedings, volume 6542 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19124-X (print), 3-642-19125-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19125-1.

Enokido:2005:EUC

[EYX^{+}05] Tomoya Enokido, Lu Yan, Bin Xiao, Daeyoung Kim, Yuanshun Dai, and Laurence T. Yang, editors. Embedded and Ubiquitous Computing: EUC 2005 Workshops: EUC 2005 Workshops: UISW, NCUS, SecUbiq, USN and TAUES, Nagasaki, Japan, December 6–9, 2005: Proceedings, volume 3823 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30803-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 E55 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30803-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3823.

Fitzgibbon:2012:CVEb

[F^{+}12a] A. W. (Andrew W.) Fitzgibbon et al., editors. Computer vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7–13, 2012, Proceedings. Part II, volume 7573 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33708-2, 3-642-33709-0 (e-book). LCCN TA1634 .E97 2012. URL http:

//www.springerlink.

com/content/978-3-642-
33708-6;   http://www.
springerlink.com/content/
978-3-642-33709-3.

Fitzgibbon:2012:CVEd

[F^{+}12b] A. W. (Andrew W.)

Fitzgibbon et al., editors. Computer vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7–13, 2012, Proceedings. Part IV, volume 7575 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33764-3, 3-642-33765-1 (e-book). LCCN TA1634 .E97 2012. URL http:

//www.springerlink.
com/content/978-3-642-
33764-2;   http://www.
springerlink.com/content/
978-3-642-33765-9.

Fitzgibbon:2012:CVEe

[F^{+}12c] A. W. (Andrew W.)

Fitzgibbon et al., editors. Computer vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7–13, 2012, Proceedings. Part V, volume 7576 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33714-7, 3-642-33715-5 (e-book). LCCN TA1634 .E97 2012. URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33714-7;   http://www.
                                        springerlink.com/content/
                                        978-3-642-33715-4.

Fitzgibbon:2012:CVEf

[F^{+}12d] A. W. (Andrew W.) Fitzgibbon et al., editors. Computer vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7–13, 2012, Proceedings. Part VI, volume 7577 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33782-1, 3-642-33783-X (e-book). LCCN TA1634 .E97 2012. URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33782-6;   http://www.
                                        springerlink.com/content/
                                        978-3-642-33783-3.

Faugeras:1990:CVE

[Fau90] O. D. Faugeras, editor. Computer vision

— ECCV 90: First European Conference on Computer Vision, Antibes, France, April 23– 27, 1990: proceedings, volume 427 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52522-X (U.S). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .E87 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0427.htm;
 http://www.springerlink.
com/content/978-0-387-
52522-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=427.

At head of title: INRIA.

Flegel:2009:DIM

[FB09] Ulrich Flegel and Danilo Bruschi, editors. Detection of Intrusions and Malware, and Vulnerability Assessment: 6th International Conference, DIMVA 2009, Como, Italy, July 9–10, 2009. Proceedings, volume 5587 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02917-5 (print), 3-642-02918-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02918-9.

Fernandez-Baca:2012:LTI

[FB12] David Fernández-Baca, editor. LATIN 2012: Theoretical Informatics: 10th Latin American Symposium, Arequipa, Peru, April 16–20, 2012. Proceedings, volume 7256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29343-3 (print), 3-642-29344-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29344-3.

Freksa:2000:SCI

[FBHW00] C. Freksa, W. Brauer, C. Habel, and K. F. Wender, editors. Spatial Cognition II: Integrating Abstract Theories, Empirical Studies, Formal Methods, and Practical Applications, volume 1849 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67584-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 420 pp. LCCN QA267.A1 L43 no.1849. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1849.htm;
 http://www.springerlink.
com/content/978-3-540-
67584-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1849.

Falcone:2003:TRS

[FBKS03] Rino Falcone, Suzanne Barber, Larry Korba, and Munindar Singh, editors. Trust, reputation, and security: theories and practice: AA-MAS 2002 International Workshop, Bologna, Italy, July 15, 2002: Selected and Invited Papers, volume 2631 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00988-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 T78 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2631.htm;
 http://www.springerlink.
com/content/978-3-540-
00988-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2631.

Proceedings of a workshop held as part of the Autonomous Agents and Multi-Agent Systems 2002 Conference (AAMAS 2002).

Fischlin:2012:PKC

[FBM12] Marc Fischlin, Johannes Buchmann, and Mark Manulis, editors. Public Key Cryptography — PKC 2012: 15th International Conference on Practice and Theory in Public Key Cryptography, Darmstadt, Germany, May 21–23. Proceedings, volume 7293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30056-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30056-1.

Falcone:2005:TAT

[FBSMS05] Rino Falcone, Suzanne Barber, Jordi Sabater-Mir, and Munindar P. Singh, editors. Trusting agents for trusting electronic societies: theory and applications in HCI and E-commerce, volume 3577 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28012-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 I5945 2003; QA76.76.I58; QA76.76.I58 I587 2003eb; QA76.76.I58 I587 2003; QA76.76.I58 I588 2003; Internet. URL

http://www.springerlink.
com/content/978-3-540-
28012-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3577.

Falcone:2008:TAS

[FBSMS08] Rino Falcone, Suzanne K.

Barber, Jordi Sabater-Mir, and Munindar P. Singh, editors. Trust in Agent Societies: 11th International Workshop, TRUST 2008, Estoril, Portugal, May 12– 13, 2008. Revised Selected and Invited Papers, volume 5396 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92802-2 (print), 3-540-92803-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92803-4.

Flammini:2011:CSR

[FBV11] Francesco Flammini, Sandro Bologna, and Valeria Vittorini, editors. Computer Safety, Reliability, and Security: 30th International Conference,SAFECOMP 2011, Naples, Italy, September 19–22, 2011. Proceedings, volume 6894 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24269-3 (print), 3-642-24270-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24270-0.

Frank:1993:SIT

[FC93] Andrew U. Frank and I. Campari, editors. Spatial information theory: a theoretical basis for GIS: European Conference, COSIT’93, Marciana Marina, Elba Island, Italy, September 19–22, 1993: proceedings, volume 716 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57207-4 (New York), 3-540-57207-4 (Berlin).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.2 .E984 1993. DM86.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0716.htm;
 http://www.springerlink.
com/content/978-0-387-
57207-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=716.

Farach-Colton:1998:CPM

[FC98] Martin Farach-Colton, editor. Combinatorial pattern matching: 9th Annual Symposium, CPM 98, Piscataway, New Jersey, USA, July 20–22, 1998: Proceedings, volume 1448 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64739-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1448.htm;
 http://www.springerlink.
com/content/978-3-540-
64739-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1448.

Farach-Colton:2004:LTI

[FC04] Martin Farach-Colton, editor. LATIN 2004: Theoretical Informatics: 6th Latin American symposium, Buenos Aires, Argentina, April 5–8, 2004: Proceedings, volume 2976 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21258-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .L348 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2976.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21258-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2976;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95852.

Fred:2004:SSS

[FCD^{+}04] Ana Fred, Terry Caelli, Robert P. W. Duin, Aurelio Campilho, and Dick de Ridder, editors. Structural, Syntactic, and Statistical Pattern Recognition: Joint IAPR International Workshops, SSPR 2004 and SPR 2004, Lisbon, Portugal, August 18–20, 2004: Proceedings, volume 3138 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22570-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 I57 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3138.
htm;   http://www.springerlink.
com/content/978-3-540-
22570-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3138;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98738.

Furon:2007:IHI

[FCDB07] Teddy Furon, François Cayre, Gwenaël Doërr, and Patrick Bas, editors. Information Hiding: 9th International Workshop, IH 2007, Saint Malo, France, June 11–13, 2007, Revised Selected Papers, volume 4567 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77369-X (print), 3-540-77370-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77370-2.

Frank:1992:TMS

[FCF92] Andrew U. Frank, I. Campari, and U. Formentini, editors. Theories and methods of spatio-temporal reasoning in geographic space, volume 639 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55966-3 (Berlin), 0-387-55966-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.2 .I575 1992.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0639.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        55966-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=639.

Freire:2004:UMN

[FCLG04] Mário Marques Freire, Prosper Chemouil, Pascal Lorenz, and Annie Gravey, editors. Universal Multiservice Networks: Third European Conference, ECUMN 2004, Porto, Portugal, October 25–27, 2004. Proceedings, volume 3262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23551-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3262.htm;
 http://www.springerlink.
com/content/978-3-540-
23551-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3262;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101785.

Fong:2002:AWB

[FCLL02] Joseph Fong, Chu Ting Cheung, Hong Va Leong, and Qing Li, editors. Advances in Web-Based Learning: First International Conference, ICWL 2002, Hong Kong, China, August 17–19, 2002. Proceedings, volume 2436 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44041-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1044.87 .I39 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2436.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44041-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2436.

Filgueiras:1993:PAI

[FD93] M. Filgueiras and Luis Damas, editors. Progress in artifical intelligence: 6th Portuguese Conference on AI, EPIA ’93, Porto, Portugal, October 6–8, 1993: proceedings, volume 727 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57287-2 (Berlin), 0-387-57287-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P67 1993.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0727.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        57287-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=727.

Favela:2003:GDI

[FD03] Jesus Favela and Dominique Decouchant, editors. Groupware: Design, Implementation, and Use: 9th International Workshop, CRIWG 2003, Autrans, France, September 28– October 2, 2003: Proceedings, volume 2806 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20117-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.43 .I684 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2806.
htm;   http://www.springerlink.
com/content/978-3-540-
20117-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2806;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13235.

Fu:2011:CCA

[FD11] Bin Fu and Ding-Zhu Du, editors. Computing and Combinatorics: 17th Annual International Conference, COCOON 2011, Dallas, TX, USA, August 14–16, 2011. Proceedings, volume 6842 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22684-1 (print), 3-642-22685-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22685-4.

Flach:2012:MLKa

[FDC12a] Peter A. Flach, Tijl De Bie, and Nello Cristianini, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2012, Bristol, UK, September 24–28, 2012. Proceedings, Part I, volume 7523 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33459-8 (print), 3-642-33460-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33460-3.

Flach:2012:MLKb

[FDC12b] Peter A. Flach, Tijl De Bie, and Nello Cristianini, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2012, Bristol, UK, September 24–28, 2012. Proceedings, Part II, volume 7524 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33485-7 (print), 3-642-33486-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33486-3.

Ficarra:2011:ANT

[FdCLJ^{+}11] Francisco V. Cipolla Ficarra, Carlos de Castro Lozano, Mauricio Pérez Jiménez, Emma Nicol, Andreas Kratky, and Miguel Cipolla-Ficarra, editors. Advances in New Technologies, Interactive Interfaces, and Communicability: First International Conference, AD-NTIIC 2010, Huerta Grande, Argentina, October 20–22, 2010, Revised Selected Papers, volume 6616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20809-6 (print), 3-642-20810-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20810-2.

Ficarra:2011:HCI

[FdCLN^{+}11] Francisco V. Cipolla Ficarra, Carlos de Castro Lozano, Emma Nicol, Andreas Kratky, and Miguel Cipolla-Ficarra, editors. Human-Computer Interaction, Tourism and Cultural Heritage: First International Workshop, HCITOCH 2010, Brescello, Italy, September 7– 8, 2010. Revised Selected Papers, volume 6529 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18347-6 (print), 3-642-18348-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18348-5.

Fischer:2010:BIR

[FDL10] Bernd Fischer, Benoît M. Dawant, and Cristian Lorenz, editors. Biomedical Image Registration: 4th International Workshop, WBIR 2010, Lübeck, Germany, July 11–13, 2010. Proceedings, volume 6204 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14365-2 (print), 3-642-14366-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14366-3.

Fraser:2012:SBS

[FdS12] Gordon Fraser and Jerffeson Teixeira de Souza, editors. Search Based Software Engineering: 4th International Symposium, SSBSE 2012, Riva del Garda, Italy, September 28–30, 2012. Proceedings, volume 7515 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33118-1 (print), 3-642-33119-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33119-0.

Federrath:2001:DPE

[Fed01] Hannes Federrath, editor. Designing privacy enhancing technologies: International Workshop on Design Issues in Anonymity and Unobservability, Berkeley, CA, USA, July 25–26, 2000: Proceedings, volume 2009 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41724-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I563 2000; QA267.A1 L43 no.2009. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2009.htm;
 http://www.springerlink.
com/content/978-3-540-
41724-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2009.

Fehr:2011:ITS

[Feh11] Serge Fehr, editor. Information Theoretic Security: 5th International Conference, ICITS 2011, Amsterdam, The Netherlands, May 21–24, 2011. Proceedings, volume 6673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20727-8 (print), 3-642-20728-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
20728-0.

Feigenbaum:1991:ACC

[Fei91] Joan Feigenbaum, editor. Advances in cryptology — CRYPTO ’91: proceedings, volume 576 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-55188-3 (New York), 3-540-55188-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0576.htm;
http://www.springerlink.
com/content/978-0-387-
55188-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=576.

Conference held Aug. 11– 15, 1991, at the University of California, Santa Barbara.

Feigenbaum:2003:DRM

[Fei03] Joan Feigenbaum, editor. Digital rights management: ACM CCS-9 workshop DRM 2002, Washington, DC, USA, November 18, 2002: Revised Papers, volume 2696 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40410-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 A345 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2696.
htm;   http://www.springerlink.
com/content/978-3-540-
40410-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2696.

Fekete:2008:AAW

[Fek08] Sándor P. Fekete, editor. Algorithmic Aspects of Wireless Sensor Networks: Fourth International Workshop, ALGOSENSORS 2008, Reykjavik, Iceland, July 2008. Revised Selected Papers, volume 5389 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92861-8 (print), 3-540-92862-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92862-1.

Fensel:2000:PSM

[Fen00] Dieter Fensel, editor. Problem-Solving Methods: Understanding, Description, Development, and Reuse, volume 1791 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67816-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 F48 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1791.htm;
 http://www.springerlink.
com/content/978-3-540-
67816-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1791.

Fermuller:1993:RMD

[Fer93] C. Fermuller. Resolution methods for the decision problem, volume 679 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56732-1 (Berlin): 0-387-56732-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 205 pp. LCCN Q335 .R454 1993. US$39.00, DM52.00. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0679.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        56732-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=679.

Ferreira:1996:PAI

[Fer96] Afonso Ferreira, editor. Parallel algorithms for irregularly structured problems: third International Workshop, IR-REGULAR ’96, Santa Barbara, CA, USA, August 19–21, 1996: proceedings, volume 1117 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61549-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .I59 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1117.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61549-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1117.

Ferreira:1998:SIS

[Fer98] Afonso Ferreira, editor. Solving irregularly structured problems in parallel: 5th International Symposium, IR-REGULAR’98, Berkeley, California, USA, August 1998: Proceedings, volume 1457 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64809-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5675 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1457.htm;
 http://www.springerlink.
com/content/978-3-540-
64809-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1457.

Ferre:2008:HPD

[Fer08] Manuel Ferre, editor.

Haptics: Perception, Devices and Scenarios: 6th International Conference, EuroHaptics 2008 Madrid, Spain, June 10– 13, 2008 Proceedings, volume 5024 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69056-5 (print), 3-540-69057-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        69057-3.

Festa:2010:EAI

[Fes10] Paola Festa, editor. Experimental Algorithms: 9th International Symposium, SEA 2010, Ischia Island, Naples, Italy, May 20–22, 2010. Proceedings, volume 6049 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13192-1 (print), 3-642-13193-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13193-6.

Ferri:2012:CTI

[FFL^{+}12] Massimo Ferri, Patrizio Frosini, Claudia Landi, Andrea Cerri, and Barbara Di Fabio, editors. Computational Topology in Image Context: 4th International Workshop, CTIC 2012, Bertinoro, Italy, May 28–30, 2012. Proceedings, volume 7309 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30237-8 (print), 3-642-30238-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30238-1.

Fischer:2005:SSC

[FFM05] Klaus Fischer, Michael Florian, and Thomas Malsch, editors. Socionics: Scalability of Complex Social Systems, volume 3413 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30707-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30707-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3413.

Feitelson:2005:JSSb

[FFRS05] Dror Feitelson, Eitan Frachtenberg, Larry Rudolph, and Uwe Schwiegelshohn, editors. Job scheduling strategies for parallel processing: 11th international workshop, JSSPP 2005, Cambridge, MA, USA, June 19, 2005: revised selected papers, volume 3834 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-31024-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .J757 2005. URL http:

//www.springerlink.
com/content/978-3-540-
31024-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3834.

Fulop:1995:ALP

[FG95] Zoltan Fulop and F. Gecseg, editors. Automata, languages and programming: 22nd international colloquium, ICALP 95, Szeged, Hungary, July 10–14, 1995: proceedings, volume 944 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60084-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.944. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0944.htm;
 http://www.springerlink.
com/content/978-3-540-
60084-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=944.

ICALP (International Colloquium on Automata, Languages, and Programming) is sponsored by the European Association for Theoretical Computer Science.

Foo:1996:PTA

[FG96] N. Y. (Norman Y.) Foo and Randy Goebel, editors. PRICAI ’96: topics in artificial intelligence: 4th Pacific Rim International Conference on Artificial Intelligence, Cairns, Australia, August 26–30, 1996: proceedings, volume 1114 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61532-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P33 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1114.htm;
http://www.springerlink.
com/content/978-3-540-
61532-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1114.

Focardi:2001:FSA

[FG01] Riccardo Focardi and Roberto Gorrieri, editors. Foundations of Security Analysis and Design: Tutorial Lectures, volume 2171 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42896-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F68 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2171.htm;
 http://www.springerlink.
com/content/978-3-540-
42896-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2171.

Lectures presented at the International School on Foundations of Security Analysis and Design, Sept. 18–30, 2000, Bertinoro, Italy.

Focardi:2004:FSA

[FG04a] Riccardo Focardi and Roberto Gorrieri, editors. Foundations of Security Analysis and Design II: FOSAD 2001/ 2002 tutorial lectures, volume 2946 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20955-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F67 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2946.
htm;   http://www.springerlink.
com/content/978-3-540-
20955-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2946;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95547.

Funk:2004:ACB

[FG04b] Peter Funk and Pedro A.

González Calero, editors. Advances in case-based reasoning: 7th European conference, ECCBR 2004, Madrid, Spain, August 30 - September 2, 2004: proceedings, volume 3155 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22882-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E33 2004. URL http:

//www.springerlink.
com/content/978-3-540-
22882-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3155;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b99702.

Flocchini:2006:SIC

[FG06] Paola Flocchini and Leszek Gasieniec, editors. Structural Information and Communication Complexity: 13th International Colloquium, SIROCCO 2006, Chester, UK, July 2– 5, 2006. Proceedings, volume 4056 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35474-3 (print), 3-540-35475-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        35475-8.

Fraser:2010:TPI

[FG10] Gordon Fraser and Angelo Gargantini, editors. Tests and Proofs: 4th International Conference, TAP 2010, Málaga, Spain, July 1–2, 2010. Proceedings, volume 6143 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13976-0 (print), 3-642-13977-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13977-2.

Fernandes:2011:ADB

[FGB11] Alvaro A. A. Fernandes, Alasdair J. G. Gray, and Khalid Belhajjame, editors. Advances in Databases: 28th British National Conference on Databases, BNCOD 28, Manchester, UK, July 12–14, 2011, Revised Selected Papers, volume 7051 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24576-5 (print), 3-642-24577-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24577-0.

Farwell:1998:MTI

[FGH98] David L. Farwell, Laurie Gerber, and Eduard H. Hovy, editors. Machine Translation and the Information Soup: Third Conference of the Association for Machine Translation in the Americas, AMTA’98, Langhorne, PA, USA, October 28– 31, 1998. Proceedings, volume 1529 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65259-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN P308.A85 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1529.htm;
 http://www.springerlink.
com/content/978-3-540-
65259-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1529.

Frappier:2010:ASM

[FGK^{+}10] Marc Frappier, Uwe Glässer, Sarfraz Khurshid, Régine Laleau, and Steve Reeves, editors. Abstract State Machines, Alloy, B and Z: Second International Conference, ABZ 2010, Orford, QC, Canada, February 22–25, 2010. Proceedings, volume 5977 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11810-0 (print), 3-642-11811-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
11811-1.

Forner:2011:MMI

[FGK^{+}11] Pamela Forner, Julio Gonzalo, Jaana Kekäläinen, Mounia Lalmas, and Marteen de Rijke, editors. Multilingual and Multimodal Information Access Evaluation: Second International Conference of the Cross-Language Evaluation Forum, CLEF 2011, Amsterdam, The Netherlands, September 19– 22, 2011. Proceedings, volume 6941 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23707-X (print), 3-642-23708-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23708-9.

Frasson:1996:ITS

[FGL96] Claude Frasson, Gilles Gauthier, and Alan M. Lesgold, editors. Intelligent tutoring systems: Third International Conference, ITS ’96 Montréal, Canada, June 12–14, 1996: proceedings, volume 1086 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61327-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.73 .I58 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1086.
htm;   http://www.springerlink.
com/content/978-3-540-
61327-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1086.

Chiefly English; some chapters in French.

Frasson:1992:ITS

[FGM92] Claude Frasson, Gilles Gauthier, and Gordon McCalla, editors. Intelligent tutoring systems: Second International Conference, ITS ’92, Montréal, Canada, June 10–12, 1992, proceedings, volume 608 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 0-387-55606-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.5 .I557 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0608.htm;

http://www.springerlink.
com/content/978-0-387-
55606-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=608.

English and French.

Freund:2008:ALT

[FGTZ08] Yoav Freund, László Györfi, György Turán, and Thomas Zeugmann, editors. Algorithmic Learning Theory: 19th International Conference, ALT 2008, Budapest, Hungary, October 13–16, 2008. Proceedings, volume 5254 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87986-2 (print), 3-540-87987-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87987-9.

Fischer-Hubner:2001:ISP

[FH01] Simone Fischer-Hübner, editor. IT-security and privacy: design and use of privacy-enhancing security mechanisms, volume 1958 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42142-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 351 pp. LCCN QA76.9.A25 F5336 2001; QA267.A1 L43 no.1958. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1958.htm;
 http://www.springerlink.
com/content/978-3-540-
42142-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1958.

Fischer-Hubner:2006:TPD

[FHFL06] Simone Fischer-Hübner, Stevel Furnell, and Costas Lambrinoudakis, editors. Trust and Privacy in Digital Business: Third International Conference, TrustBus 2006, Kraków, Poland, September 4–8, 2006. Proceedings, volume 4083 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37750-6 (print), 3-540-37752-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37752-8.

Fischer-Hubner:2011:PET

[FHH11] Simone Fischer-Hübner and Nicholas Hopper, editors. Privacy Enhancing Technologies: 11th International Symposium, PETS 2011, Waterloo, ON, Canada, July 27– 29, 2011. Proceedings, volume 6794 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22262-5 (print), 3-642-22263-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22263-4.

Fagerholm:2002:APC

[FHJ^{+}02] Juha Fagerholm, Juha Haataja, Jari Järvinen, Mikko Lyly, Peter Råback, and Ville Savolainen, editors. Applied Parallel Computing. Advanced Scientific Computing: 6th International Conference, PARA 2002, Espoo, Finland, June 15–18, 2002. Proceedings, volume 2367 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43786-X (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2367.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43786-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2367.

Fischer-Hubner:2012:TPS

[FHKQ12] Simone Fischer-Hübner, Sokratis Katsikas, and Gerald Quirchmayr, editors. Trust, Privacy and Security in Digital Business: 9th International Conference, TrustBus 2012, Vienna, Austria, September 3–7, 2012. Proceedings, volume 7449 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32286-7 (print), 3-642-32287-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32287-7.

Fischer-Hubner:2009:TPS

[FHLP09] Simone Fischer-Hübner, Costas Lambrinoudakis, and Günther Pernul, editors. Trust, Privacy and Security in Digital Business: 6th International Conference, Trust-Bus 2009, Linz, Austria, September 3–4, 2009. Proceedings, volume 5695 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03747-X (print), 3-642-03748-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03748-1.

Flynn:1982:MSD

[FHM82] M. J. Flynn, N. R. Harris, and D. P. McCarthy, editors. Microcomputer system design: an advanced course, Trinity College, Dublin, June 1981, volume 126 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11172-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 .M5. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0126.
htm;   http://www.springerlink.
com/content/978-0-387-
11172-8;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=126.

Fossorier:2003:AAA

[FHP03] Marc Fossorier, Tom Høholdt, and Alain Poli, editors. Applied algebra, algebraic algorithms, and errorcorrecting codes: 15th International Symposium, AAECC-15, Toulouse, France, May 12–16, 2003: Proceedings, volume 2643 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40111-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2643.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40111-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2643.

Also available via the

World Wide Web.

Frieze:2011:AMW

[FHP11] Alan Frieze, Paul Horn, and Paweł Prałat, editors. Algorithms and Models for the Web Graph: 8th International Workshop, WAW 2011, Atlanta, GA, USA, May 27–29, 2011. Proceedings, volume 6732 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21285-9 (print), 3-642-21286-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21286-4.

Fiadeiro:2005:ACC

[FHRR05] José Luiz Fiadeiro, Neil Harman, Markus Roggenbach, and Jan Rutten, editors. Algebra and coalgebra in computer science: first international conference, CALCO 2005, Swansea, UK, September 3–6, 2005: proceedings, volume 3629 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28620-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 C33 2005. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3629;   http:

                                        //www.loc.gov/catdir/
                                        toc/fy0604/2005931121.
                                        html;   http://www.
                                        springerlink.com/content/
                                        978-3-540-28620-2.

Fitzgerald:2005:FFM

[FHT05] John Fitzgerald, Ian Hayes, and Andrzej Tarlecki, editors. FM 2005: Formal Methods: International Symposium of Formal Methods Europe, Newcastle, UK, July 18–22, 2005: Proceedings, volume 3582 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27882-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I593 2005.

                                        URL http://springerlink.
                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3582;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        27882-5.

Freksa:1998:SCI

[FHW98] Christian Freksa, Christopher Habel, and Karl Friedrich Wender, editors. Spatial Cognition: An Interdisciplinary Approach to Representing and Processing Spatial Knowledge, volume 1404 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64603-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 489 pp. LCCN QA267.A1 L43 no.1404.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1404.
htm;   http://www.springerlink.
com/content/978-3-540-
64603-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1404.

Franklin:2008:CNS

[FHW08] Matthew K. Franklin, Lucas Chi Kwong Hui, and Duncan S. Wong, editors. Cryptology and Network Security: 7th International Conference, CANS 2008, Hong-Kong, China, December 2–4, 2008. Proceedings, volume 5339 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89640-6 (print), 3-540-89641-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        89641-8.

Fischer-Hubner:2012:PET

[FHW12] Simone Fischer-Hübner and Matthew Wright, editors. Privacy Enhancing Technologies: 12th International Symposium, PETS 2012, Vigo, Spain, July 11–13, 2012. Proceedings, volume 7384 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31679-4 (print), 3-642-31680-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31680-7.

Fitzgerald:2008:TAC

[FHY08] John S. Fitzgerald, Anne E. Haxthausen, and Husnu Yenigun, editors. Theoretical Aspects of Computing — ICTAC 2008: 5th International Colloquium, Istanbul, Turkey, September 1–3, 2008. Proceedings, volume 5160 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85761-3 (print), 3-540-85762-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85762-4.

Fiadeiro:2008:FAS

[FI08] José Luiz Fiadeiro and Paola Inverardi, editors. Fundamental Approaches to Software Engineering: 11th International Conference, FASE 2008, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2008, Budapest, Hungary, March 29–April 6, 2008. Proceedings, volume 4961 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78742-9 (print), 3-540-78743-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78743-3.

Fiadeiro:1999:RTA

[Fia99] Jose Luiz Fiadeiro, editor. Recent trends in algebraic development techniques: 13th International Workshop, WADT’98, Lisbon, Portugal, April 2–4, 1998: Selected Papers, volume 1589 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66246-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 W33 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1589.htm;
 http://www.springerlink.
com/content/978-3-540-
66246-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1589.

Ferri:2001:APR

[FIAP01] Francesc Ferri, José M. Iñesta, Adnan Amin, and Pavel Pudil, editors. Advances in pattern recognition: joint IAPR International Workshops SSPR’2000 and SPR 2000, Alicante, Spain, August 30–September 1, 2000: Proceedings, volume 1876 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-67946-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 I59 2000.

URL http://link.springer-
ny.com/link/service/

series/0558/tocs/t1876.
htm;   http://www.springerlink.
com/content/978-3-540-
67946-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1876.

Fich:2003:DCI

[Fic03] Faith Ellen Fich, editor. Distributed Computing: 17th International Conference, DISC 2003, Sorrento, Italy, October 1–3, 2002: Proceedings, volume 2848 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20184-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D438 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2848.
htm;   http://www.springerlink.
com/content/978-3-540-
20184-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2848;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13831.

Field:2002:CPE

[Fie02] Anthony J. Field, editor.

Computer performance evaluation: modelling techniques and tools:

12th International Conference, TOOLS 2002, London, UK, April 14– 17, 2002: Proceedings, volume 2324 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43539-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.E94 I55 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2324.htm;
 http://www.springerlink.
com/content/978-3-540-
43539-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2324.

Filgueiras:1991:NLP

[Fil91] M. Filgueiras, editor. Natural language processing: EAIA ’90, 2nd Advanced School in Artificial Intelligence, Guarda, Portugal, October 8–12, 1990: proceedings, volume 476 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-53678-7 (New York), 3-540-53678-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A34 1990. DM42.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0476.htm;
http://www.springerlink.
com/content/978-0-387-
53678-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=476.

Fossorier:2006:AAA

[FILP06] Marc Fossorier, Hideki Imai, Shu Lin, and Alain Poli, editors. Applied Algebra, Algebraic Algorithms and Error-Correcting Codes: 16th International Symposium, AAECC-16, Las Vegas, NV, USA, February 20–24, 2006. Proceedings, volume 3857 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31423-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31423-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3857.

Finance:1999:FAS

[Fin99] Jean Pierre Finance, editor. Fundamental approaches to software engineering: Second International Conference, FASE’99, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS’99, Amsterdam, The Netherlands, March 22–28, 1999: Proceedings, volume 1577 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65718-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F37 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1577.htm;
 http://www.springerlink.
com/content/978-3-540-
65718-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1577.

Fisher:2008:FMS

[Fis08] Jasmin Fisher, editor. Formal Methods in Systems Biology: First International Workshop, FMSB 2008, Cambridge, UK, June 4–5, 2008. Proceedings, volume 5054 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68410-7 (print), 3-540-68413-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68413-8.

Fischlin:2009:TCC

[Fis09] Marc Fischlin, editor.

Topics in Cryptology — CT-RSA 2009: The Cryptographers’ Track at the RSA Conference 2009, San Francisco, CA, USA, April 20–24, 2009. Proceedings, volume 5473 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00861-5 (print), 3-642-00862-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00862-7.

Fitch:1984:EIS

[Fit84] John Fitch, editor. EU-ROSAM 84: International Symposium on Symbolic and Algebraic Computation, Cambridge, England, July 9–11, 1984, volume 174 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CODEN LNCSD9. ISBN 0-387-13350-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E88 1984. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0174.htm;
 http://www.springerlink.
com/content/978-0-387-
13350-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=174.

Fitch:1993:DIS

[Fit93] John Fitch, editor. Design and implementation of symbolic computation systems: International Symposium, DISCO ’92, Bath, UK, April 13–15, 1992: proceedings, volume 721 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57272-4 (Berlin), 0-387-57272-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0721.htm;
 http://www.springerlink.
com/content/978-3-540-
57272-5;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=721.

Finkel:1992:SAS

[FJ92] A. Finkel and Matthias Jantzen, editors. STACS 92: 9th Annual Symposium on Theoretical Aspects of Computer Science, Cachan, France, February 13–15, 1992: proceedings, volume 577 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55210-3 (Berlin), 0-387-55210-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.577. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0577.htm;
http://www.springerlink.
com/content/978-3-540-
55210-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=577.

Foresti:2010:DAS

[FJ10] Sara Foresti and Sushil Jajodia, editors. Data and Applications Security and Privacy XXIV: 24th Annual IFIP WG 11.3 Working Conference, Rome, Italy, June 21–23, 2010. Proceedings, volume 6166 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13738-5 (print), 3-642-13739-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13739-6.

Fitzgerald:1997:FIA

[FJL97] John Fitzgerald, C. B. Jones, and Peter Lucas, editors. FME ’97, industrial applications and strengthened foundations of formal methods: 4th International Symposium of Formal Methods Europe, Graz, Austria, September 15–19, 1997: proceedings, volume 1313 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63533-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I577 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1313.htm;
 http://www.springerlink.
com/content/978-3-540-
63533-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1313.

Futatsugi:2006:AMC

[FJM06] Kokichi Futatsugi, Jean-Pierre Jouannaud, and José Meseguer, editors. Algebra, Meaning, and Computation: Essays dedicated to Joseph A. Goguen on the Occasion of His 65th Birthday, volume 4060 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35462-X (print), 3-540-35464-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35464-2.

Fasel:1987:GRP

[FK87] Joseph H. Fasel and Robert M. Keller, editors. Graph reduction: proceedings of a workshop, Santa Fe, New Mexico, USA, September 29–October 1, 1986, volume 279 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-18420-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A23 G7 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0279.htm;
 http://www.springerlink.
com/content/978-0-387-
18420-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=279.

“Sponsored by the Los Alamos National Laboratory of the University of California and Microelectronics and Computer Technology Corporation (MCC)”–Pref.

Frank:1995:SIT

[FK95a] Andrew U. Frank and Werner Kuhn, editors. Spatial information theory: a theoretical basis for GIS: international conference, COSIT’95, Semmering, Austria, September 21–23, 1995: proceedings, volume 988 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60392-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.2 .C68 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0988.
htm;   http://www.springerlink.
com/content/978-3-540-
60392-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=988.

Froidevaux:1995:SQA

[FK95b]  Christine  Froidevaux

and Jurg Kohlag, editors. Symbolic and quantitative approaches to reasoning and uncertainty: European Conference ECSQARU ’95, Fribourg, Switzerland, July 1995: proceedings, volume 946 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60112-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.946.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0946.
htm;   http://www.springerlink.
com/content/978-3-540-
60112-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=946.

Fieker:2002:ANT

[FK02] Claus Fieker and David R.

Kohel, editors. Algorithmic Number Theory: 5th International Symposium, ANTS-V, Sydney, Australia, July 7– 12, 2002. Proceedings, volume 2369 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43863-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA241 .A432 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2369.htm;
 http://www.springerlink.
com/content/978-3-540-
43863-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2369.

Fuller:2009:MEL

[FK09] Richard Fuller and Xenofon D. Koutsoukos, editors. Mobile Entity Localization and Tracking in GPS-less Environnments: Second International Workshop, MELT 2009, Orlando, FL, USA, September 30, 2009. Proceedings, volume 5801 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04378-X (print), 3-642-04385-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
04385-7.

Farzindar:2010:AAI

[FK10] Atefeh Farzindar and Vlado Kešelj, editors. Advances in Artificial Intelligence: 23rd Canadian Conference on Artificial Intelligence, Canadian AI 2010, Ottawa, Canada, May 31 — June 2, 2010. Proceedings, volume 6085 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13058-5 (print), 3-642-13059-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13059-5.

Flanagan:2012:TAC

[FK12a] Cormac Flanagan and Barbara König, editors. Tools and Algorithms for the Construction and Analysis of Systems: 18th International Conference, TACAS 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28755-7 (print), 3-642-28756-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28756-5.

Fomin:2012:ATS

[FK12b] Fedor V. Fomin and Petteri Kaski, editors. Algorithm Theory — SWAT 2012: 13th Scandinavian Symposium and Workshops, Helsinki, Finland, July 4–6, 2012. Proceedings, volume 7357 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31154-7 (print), 3-642-31155-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31155-0.

Feridun:2002:MTC

[FKB02] Metin Feridun, Peter Kropf, and Gilbert Babin, editors. Management Technologies for E-Commerce and E-Business Applications:

13th IFIP/IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2002, Montréal, Canada, October 21–23, 2002. Proceedings, volume 2506 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00080-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I33835 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2506.htm;
 http://www.springerlink.
com/content/978-3-540-
00080-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2506.

France:2012:MDE

[FKBA12] Robert B. France, Jürgen Kazmeier, Ruth Breu, and Colin Atkinson, editors. Model Driven Engineering Languages and Systems: 15th International Conference, MOD-ELS 2012, Innsbruck, Austria, September 30– October 5, 2012. Proceedings, volume 7590 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33665-5 (print), 3-642-33666-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33666-9.

France:2011:MFA

[FKBP11] Robert B. France, Jochen M. Kuester, Behzad Bordbar, and Richard F. Paige, editors. Modelling Foundations and Applications: 7th European Conference, ECMFA 2011, Birmingham, UK, June 6 — 9, 2011 Proceedings, volume 6698 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21469-X (print), 3-642-21470-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21470-7.

Fiala:2004:MFC

[FKK04] Jirà Fiala, Václav Koubek, and Jan Kratochvàl, editors. Mathematical Foundations of Computer Science 2004: 20th International Symposium, MFCS 2004, Prague, Czech Republic, August 22–27, 2004. Proceedings, volume 3153 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22823-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 S96 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3153.
htm;   http://www.springerlink.
com/content/978-3-540-
22823-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3153;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99679.

Feldman:2009:NGI

[FKK09] Yishai A. Feldman, Donald Kraft, and Tsvi Kuflik, editors. Next Generation Information Technologies and Systems: 7th International Conference, NGITS 2009, Haifa, Israel, June 16– 18, 2009. Revised Selected Papers, volume 5831 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04940-0 (print), 3-642-04941-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04941-5.

Freksa:2005:SCI

[FKKB^{+}05] Christian Freksa, Markus Knauff, Bernd Krieg-Brückner, Bernhard Nebel, and Thomas Barkowsky, editors. Spatial cognition IV: reasoning, action, interaction: international conference on spatial cognition 2004, Frauenchiemsee, Germany, October 11–13, 2004. Revised selected papers, volume 3343 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25048-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387 .S74 2004.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        25048-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3343;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106616.

Furnell:2008:TPS

[FKL08] Steven Furnell, Sokratis K.

Katsikas, and Antonio Lioy, editors. Trust, Privacy and Security in Digital Business: 5th International Conference, TrustBus 2008 Turin, Italy, September 4–5, 2008 Proceedings, volume 5185 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85734-6 (print), 3-540-85735-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85735-8.

Fuhr:2008:FAX

[FKLT08] Norbert Fuhr, Jaap Kamps, Mounia Lalmas, and Andrew Trotman, editors. Focused Access to XML Documents: 6th International Workshop of the Initiative for the Evaluation of XML Retrieval, INEX 2007 Dagstuhl Castle, Germany, December 17– 19, 2007. Selected Papers, volume 4862 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85901-2 (print), 3-540-85902-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85902-4.

Fyfe:2008:IDE

[FKLY08] Colin Fyfe, Dongsup Kim, Soo-Young Lee, and Hujun Yin, editors. Intelligent Data Engineering and Automated Learning — IDEAL 2008: 9th International Conference Daejeon, South Korea, November 2–5, 2008 Proceedings, volume 5326 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88905-1 (print), 3-540-88906-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88906-9.

Franconi:2007:SWR

[FKM07] Enrico Franconi, Michael Kifer, and Wolfgang May, editors. The Semantic Web: Research and Applications: 4th European Semantic Web Conference, ESWC 2007, Innsbruck, Austria, June 3–7, 2007. Proceedings, volume 4519 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72666-7 (print), 3-540-72667-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72667-8.

Freire:2008:PAD

[FKM08] Juliana Freire, David Koop, and Luc Moreau, editors. Provenance and Annotation of Data and Processes: Second International Provenance and Annotation Workshop, IPAW 2008, Salt Lake City, UT, USA, June 17–18, 2008. Revised Selected Papers, volume 5272 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89964-2 (print), 3-540-89965-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89965-5.

Fiala:2009:CAI

[FKM09] Jiří Fiala, Jan Kratochvíl, and Mirka Miller, editors. Combinatorial Algorithms: 20th International Workshop, IWOCA 2009, Hradec nad Moravicí, Czech Republic, June 28–July 2, 2009, Revised Selected Papers, volume 5874 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10216-6 (print), 3-642-10217-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10217-2.

Felici:1999:CSR

[FKP99] Massimo Felici, Karama Kanoun, and Alberto Pasquini, editors. Computer safety, reliability and security: 18th International Conference, SAFECOMP’99, Toulouse, France, September 27–29, 1999: Proceedings, volume 1698 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66488-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 I55 1999. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t1698.htm;
http://www.springerlink.
com/content/978-3-540-
66488-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1698.

Famili:2005:AID

[FKP^{+}05] A. Fazel Famili, Joost N.

Kok, José M. Peña, Arno Siebes, and Ad Feelders, editors. Advances in intelligent data analysis VI: 6th international symposium on intelligent data analysis, IDA 2005, Madrid, Spain, September 8–10, 2005, proceedings, volume 3646 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28795-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28795-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3646.

Fernstrom:1986:LAA

[FKS86] Christer Fernstrom, Ivan Kruzela, and Bertil Svensson. LUCAS associative array processor: design, programming, and application studies, volume 216 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16445-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 323 pp. LCCN QA76.5 .F37351 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0216.htm;
 http://www.springerlink.
com/content/978-0-387-
16445-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=216.

Freksa:2007:KAA

[FKS07] Christian Freksa, Michael Kohlhase, and Kerstin Schill, editors. KI 2006: Advances in Artificial Intelligence: 29th Annual German Conference on AI, KI 2006, Bremen, Germany, June 14–17, 2006. Proceedings, volume 4314 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69911-2 (print), 3-540-69912-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
69912-5.

Floreen:2010:PCI

[FKS10] Patrik Floréen, Antonio Krüger, and Mirjana Spasojevic, editors. Pervasive Computing: 8th International Conference, Pervasive 2010, Helsinki, Finland, May 17–20, 2010. Proceedings, volume 6030 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12653-7 (print), 3-642-12654-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12654-3.

Fong:2008:HLE

[FKW08] Joseph Fong, Reggie Kwan, and Fu Lee Wang, editors. Hybrid Learning and Education: First International Conference, ICHL 2008 Hong Kong, China, August 13–15, 2008 Proceedings, volume 5169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85169-0 (print), 3-540-85170-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85170-7.

Falsafi:2000:NBP

[FL00] Babak Falsafi and Mario Lauria, editors. Networkbased parallel computing: communication, architecture, and applications: 4th International Workshop, CANPC 2000, Toulouse, France, January 8, 2000: Proceedings, volume 1797 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67879-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .C374 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1797.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67879-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1797.

Ferragina:2008:CPM

[FL08] Paolo Ferragina and Gad M. Landau, editors. Combinatorial Pattern Matching: 19th Annual Symposium, CPM 2008, Pisa, Italy, June 18–20, 2008 Proceedings, volume 5029 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69066-2 (print), 3-540-69068-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69068-9.

Feuerlicht:2009:SOC

[FL09] George Feuerlicht and Winfried Lamersdorf, editors. Service-Oriented Computing — ICSOC 2008 Workshops: IC-SOC 2008 International Workshops, Sydney, Australia, December 1st, 2008, Revised Selected Papers, volume 5472 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01246-9 (print), 3-642-01247-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01247-1.

Frasconi:2011:ILP

[FL11] Paolo Frasconi and Francesca A.

Lisi, editors. Inductive Logic Programming: 20th International Conference, ILP 2010, Florence, Italy, June 27– 30, 2010. Revised Papers, volume 6489 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21294-8 (print), 3-642-21295-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21295-6.

Flener:1999:LBP

[Fle99] Pierre Flener, editor. Logic-based program synthesis and transformation: 8th International Workshop, LOPSTR’98, Manchester, UK, June 15–19, 1998: Selected Papers, volume 1559 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65765-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1559.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1559.
htm;   http://www.springerlink.
com/content/978-3-540-
65765-1;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1559.

Flesca:2002:LAI

[Fle02] Sergio Flesca, editor.

Logics in artificial intelligence: 8th European Conference, JELIA 2002, Cosenza, Italy, September 23–26, 2002: Proceedings, volume 2424 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44190-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .E985 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2424.htm;
http://www.springerlink.
com/content/978-3-540-
44190-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2424.

Also available via the World Wide Web.

Floerkemeier:2008:ITF

[FLF^{+}08] Christian Floerkemeier, Marc Langheinrich, Elgar Fleisch, Friedemann Mattern, and Sanjay E. Sarma, editors. The Internet of Things: First International Conference, IOT 2008, Zurich, Switzerland, March 26– 28, 2008. Proceedings, volume 4952 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78730-5 (print), 3-540-78731-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78731-0.

Freire:2003:HSN

[FLL03] Mario Marques Freire, Pascal Lorenz, and Mike Myung-Ok Lee, editors. High-speed Networks and Multimedia Communications: 6th IEEE International Conference, HSNMC 2003, Estoril, Portugal, July 23– 25, 2003: Proceedings, volume 2720 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40542-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .H75 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2720.htm;
 http://www.springerlink.
com/content/978-3-540-
40542-9;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2720.

Ferreira:2010:PPP

[FLMG10] Fernando Ferreira, Benedikt Löwe, Elvira Mayordomo, and Luís Mendes Gomes, editors. Programs, Proofs, Processes: 6th Conference on Computability in Europe, CiE 2010, Ponta Delgada, Azores, Portugal, June 30 — July 4, 2010. Proceedings, volume 6158 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13961-2 (print), 3-642-13962-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13962-8.

Fuhr:2006:AXI

[FLMK06] Norbert Fuhr, Mounia Lalmas, Saadia Malik, and Gabriella Kazai, editors. Advances in XML Information Retrieval and Evaluation: 4th International Workshop of the Initiative for the Evaluation of XML Retrieval, INEX 2005, Dagstuhl Castle, Germany, November 28–30, 2005. Revised Selected Papers, volume 3977 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34962-6 (print), 3-540-34963-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34963-1.

Fuhr:2005:AXI

[FLMS05] Norbert Fuhr, Mounia Lalmas, Saadia Malik, and Zoltán Szlávik, editors. Advances in XML information retrieval: third international workshop of the initiative for the evaluation of XML retrieval, INEX 2004, Dagstuhl Castle, Germany, December 6–8, 2004. Revised selected papers, volume 3493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26166-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26166-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3493;
 http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b136250.

Furnell:2011:TPS

[FLP11] Steven Furnell, Costas Lambrinoudakis, and Günther Pernul, editors. Trust, Privacy and Security in Digital Business: 8th International Conference, TrustBus 2011, Toulouse, France, August 29 — September 2, 2011. Proceedings, volume 6863 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22889-5 (print), 3-642-22890-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22890-2.

Finkel:2012:RPI

[FLP12a] Alain Finkel, Jérôme Leroux, and Igor Potapov, editors. Reachability Problems: 6th International Workshop, RP 2012, Bordeaux, France, September 17–19, 2012. Proceedings, volume 7550 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33511-X (print), 3-642-33512-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33512-9.

Fitzgibbon:2012:CVEa

[FLP^{+}12b] Andrew Fitzgibbon, Svetlana Lazebnik, Pietro Perona, Yoichi Sato, and Cordelia Schmid, editors. Computer Vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7– 13, 2012, Proceedings, Part I, volume 7572 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33717-1 (print), 3-642-33718-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33718-5.

Fitzgibbon:2012:CVEc

[FLP^{+}12c] Andrew Fitzgibbon, Svetlana Lazebnik, Pietro Perona, Yoichi Sato, and Cordelia Schmid, editors. Computer Vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7– 13, 2012, Proceedings, Part III, volume 7574 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33711-2 (print), 3-642-33712-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33712-3.

Fitzgibbon:2012:CVEg

[FLP^{+}12d] Andrew Fitzgibbon, Svetlana Lazebnik, Pietro Perona, Yoichi Sato, and Cordelia Schmid, editors. Computer Vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7– 13, 2012, Proceedings, Part VII, volume 7578 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33785-6 (print), 3-642-33786-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33786-4.

Frey:2011:AHM

[FLR11] Hannes Frey, Xu Li, and Stefan Ruehrup, editors. Ad-hoc, Mobile, and Wireless Networks: 10th International Conference, ADHOC-NOW 2011, Paderborn, Germany, July 18–20, 2011. Proceedings, volume 6811 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22449-0 (print), 3-642-22450-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22450-8.

Fuks:2005:GDI

[FLS05] Hugo Fuks, Stephan Lukosch, and Ana Carolina Salgado, editors. Groupware: design, implementation, and use: 11th international workshop, CRIWG 2005, Porto de Galinhas, Brazil, September 25–29, 2005: proceedings, volume 3706 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29110-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.43 .I694 2005.

URL http://www.springerlink.
com/content/978-3-540-

29110-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3706.

Farre:2006:IAA

[FLS06] Jacques Farré, Igor Litovsky, and Sylvain Schmitz, editors. Implementation and application of automata: 10th international conference, CIAA 2005, Sophia Antipolis, France, June 27– 29, 2005, revised selected papers, volume 3845 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31023-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31023-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3845.

Fuhr:2007:CEX

[FLT07] Norbert Fuhr, Mounia Lalmas, and Andrew Trotman, editors. Comparative Evaluation of XML Information Retrieval Systems: 5th International Workshop of the Initiative for the Evaluation of XML Retrieval, INEX 2006, Dagstuhl Castle, Germany, December 17–20, 2006, Revised and Selected Papers, volume 4518 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73887-8 (print), 3-540-73888-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73888-6.

Forbrig:2002:ISD

[FLUV02] Peter Forbrig, Quentin Limbourg, Bodo Urban, and Jean Vanderdonckt, editors. Interactive Systems. Design, Specification, and Verification: 9th International Workshop, DSV-IS 2002 Rostock, Germany, June 12– 14, 2002. Revised Papers, volume 2545 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00266-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.I58 D78 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2545.htm;
 http://www.springerlink.

com/content/978-3-540-
00266-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2545.

Also available via the World Wide Web.

Fernandes:2011:GTT

[FLVS11] João M. Fernandes, Ralf Lämmel, Joost Visser, and João Saraiva, editors. Generative and Transformational Techniques in Software Engineering III: International Summer School, GTTSE 2009, Braga, Portugal, July 6–11, 2009. Revised Papers, volume 6491 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18022-1 (print), 3-642-18023-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18023-1.

Feng:2006:ISC

[FLY06] Dengguo Feng, Dongdai Lin, and Moti Yung, editors. Information Security and Cryptology: First SKLOIS Conference, CISC 2005, Beijing, China, December 15–17, 2005. Proceedings, volume 3822 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30855-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30855-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3822.

Fontet:1984:SST

[FM84] M. Fontet and Kurt Mehlhorn, editors. STACS 84: Symposium of Theoretical Aspects of Computer Science, Paris

[April] 11–13, 1984, volume 166 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-12920-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S958 1984. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0166.htm;
 http://www.springerlink.
com/content/978-0-387-
12920-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=166.

Fuhrmann:1991:LTC

[FM91] A. Fuhrmann and M. Morreau, editors. The Logic of theory change: workshop, Konstanz, FRG, October 13–15, 1989, proceedings, volume 465 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-53567-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .L5663 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0465.htm;
http://www.springerlink.
com/content/978-0-387-
53567-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=465.

Futatsugi:1996:OTA

[FM96] Kokichi Futatsugi and Satoshi Matsuoka, editors. Object technologies for advanced software: Second JSSST International Symposium, ISOTAS ’96, Kanazawa, Japan, March 11–15, 1996: proceedings, volume 1049 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60954-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .I55 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1049.htm;
 http://www.springerlink.
com/content/978-3-540-
60954-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1049.

Fdida:1997:MAS

[FM97] Serge Fdida and Michele Morganti, editors. Multimedia applications, services, and techniques

— ECMAST ’97: second European Conference, Milan, Italy, May 21–23, 1997: proceedings, volume 1242 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63078-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .E25 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1242.htm;
 http://www.springerlink.
com/content/978-3-540-
63078-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1242.

Freksa:1999:SIT

[FM99] C. Freksa and David M.

Mark, editors. Spatial information theory: cognitive and computational foundations of geographic information science: International Conference COSIT ’99 Stade, Germany, August 25–29, 1999: Proceedings, volume 1661 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66365-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.212 .C68 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1661.htm;
http://www.springerlink.
com/content/978-3-540-
66365-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1661.

Ferscha:2004:PCS

[FM04] Alois Ferscha and Friedemann Mattern, editors. Pervasive Computing: Second International Conference, PER-VASIVE 2004, Linz/ Vienna, Austria, April 21–23, 2004. Proceedings, volume 3001 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21835-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3001.htm;
 http://www.springerlink.
com/content/978-3-540-
21835-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3001;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96922.

Favela:2004:AWI

[FMC04] Jesus Favela, Ernestina Menasalvas, and Edgar Chávez, editors. Advances in Web Intelligence: Second International Atlantic Web Intelligence Conference, AWIC 2004, Cancun, Mexico, May 16–19, 2004: Proceedings, volume 3034 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22009-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .I5726 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3034.
htm;   http://www.springerlink.
com/content/978-3-540-
22009-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3034;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97883.

Fusiello:2012:CVEa

[FMC12a] Andrea Fusiello, Vittorio Murino, and Rita Cucchiara, editors. Computer Vision — ECCV 2012. Workshops and Demonstrations: Florence, Italy, October 7– 13, 2012, Proceedings, Part I, volume 7583 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33862-3 (print), 3-642-33863-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33863-2.

Fusiello:2012:CVEb

[FMC12b] Andrea Fusiello, Vittorio Murino, and Rita Cucchiara, editors. Computer Vision — ECCV 2012. Workshops and Demonstrations: Florence, Italy, October 7– 13, 2012, Proceedings, Part II, volume 7584 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33867-4 (print), 3-642-33868-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33868-7.

Fusiello:2012:CVEc

[FMC12c] Andrea Fusiello, Vittorio Murino, and Rita Cucchiara, editors. Computer Vision — ECCV 2012. Workshops and Demonstrations: Florence, Italy, October 7– 13, 2012, Proceedings, Part III, volume 7585 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33884-4 (print), 3-642-33885-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33885-4.

Forsyth:1999:SCG

[FMdGC99] David Forsyth, Joseph L.

Mundy, Vito di Gesú, and Roberto Cipolla, editors. Shape, contour, and grouping in computer vision: Proceedings of an International Workshop, volume 1681 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66722-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .S47 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1681.htm;
 http://www.springerlink.
com/content/978-3-540-
66722-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1681.

Floreano:1999:AAL

[FMN99] Dario Floreano, Francesco Mondada, and J. D. (Jean-Daniel) Nicoud, editors. Advances in artificial life: 5th European Conference, ECAL’99, Lausanne, Switzerland, September, 1999: Proceedings, volume 1674 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66452-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .E87 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1674.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66452-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1674.

Franke:2006:PRD

[FMNS06] Katrin Franke, Klaus-Robert Müller, Bertram Nickolay, and Ralf Schäfer, editors. Pattern Recognition: 28th DAGM Symposium, Berlin, Germany, September 12– 14, 2006. Proceedings, volume 4174 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44412-2 (print), 3-540-44414-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        44414-5.

Fiadeiro:2005:RTA

[FMO05] José Luiz Fiadeiro, Peter D. Mosses, and Fernando Orejas, editors. Recent trends in algebraic development techniques: 17th international workshop, WADT 2004, Barcelona, Spain, March 27–29, 2004: Revised selected papers, volume 3423 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25327-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 W33 2004.

URL http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3423;
 http://www.springerlink.
com/content/978-3-540-
25327-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3423;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107129.

Fichtinger:2011:MICa

[FMP11a] Gabor Fichtinger, Anne Martel, and Terry Peters, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2011: 14th International Conference, Toronto, Canada, September 18– 22, 2011, Proceedings, Part I, volume 6891 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23622-7 (print), 3-642-23623-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23623-5.

Fichtinger:2011:MICb

[FMP11b] Gabor Fichtinger, Anne Martel, and Terry Peters, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2011: 14th International Conference, Toronto, Canada, September 18– 22, 2011, Proceedings, Part II, volume 6892 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23628-6 (print), 3-642-23629-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23629-7.

Fichtinger:2011:MICc

[FMP11c] Gabor Fichtinger, Anne Martel, and Terry Peters, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2011: 14th International Conference, Toronto, Canada, September 18– 22, 2011, Proceedings, Part III, volume 6893 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23625-1 (print), 3-642-23626-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23626-6.

Francesconi:2010:SPL

[FMPT10] Enrico Francesconi, Simonetta Montemagni, Wim Peters, and Daniela Tiscornia, editors. Semantic Processing of Legal Texts: Where the Language of Law Meets the Law of Language, volume 6036 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12836-X (print), 3-642-12837-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12837-0.

Frid:2009:CST

[FMRW09] Anna Frid, Andrey Morozov, Andrey Rybalchenko, and Klaus W. Wagner, editors. Computer Science — Theory and Applications: Fourth International Computer Science Symposium in Russia, CSR 2009, Novosibirsk, Russia, August 18–23, 2009. Proceedings, volume 5675 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03350-4 (print), 3-642-03351-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03351-3.

Fleischer:2002:EAA

[FMS02] Rudolf Fleischer, Bernard M. E. Moret, and Erik Meineche Schmidt, editors. Experimental Algorithmics: From Algorithm Design to Robust and Efficient Software, volume 2547 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00346-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 S436 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2547.htm;
 http://www.springerlink.
com/content/978-3-540-
00346-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2547.

Papers from the first Schloss Dagstuhl Seminar on Experimental Algorithmics, held in September 2000. Also available via the World Wide Web.

Ferreira:2010:ABC

[FMS10] Carlos E. Ferreira, Satoru Miyano, and Peter F. Stadler, editors. Advances in Bioinformatics and Computational Biology: 5th Brazilian Symposium on Bioinformatics, BSB 2010, Rio de Janeiro, Brazil, August 31–September 3, 2010. Proceedings, volume 6268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15059-4 (print), 3-642-15060-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        15060-9.

Fiege:2001:ECS

[FMW01] Ludger Fiege, Gero Mühl, and Uwe Wilhelm, editors. Electronic Commerce: Second International Workshop, WELCOM 2001 Heidelberg, Germany, November 16–17, 2001. Proceedings, volume 2232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42878-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.32 .W464 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2232.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42878-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2232.

Futatsugi:2004:SST

[FMY04] Kokichi Futatsugi, Fumio Mizoguchi, and Naoki Yonezaki, editors. Software Security— Theories and Systems: Second MEXT-NSF-JSPS International Symposium, ISSS 2003, Tokyo, Japan, November 4– 6, 2003. Revised Papers, volume 3233 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23635-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3233.htm;
 http://www.springerlink.
com/content/978-3-540-
23635-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3233;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b102118.

Ferretti:2005:DCI

[FMZ05] Claudio Ferretti, Giancarlo Mauri, and Claudio Zandron, editors. DNA Computing: 10th International Workshop on DNA Computing, DNA10, Milan, Italy, June 7–10, 2004, Revised Selected Papers, volume 3384 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26174-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26174-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3384.

Freitas:2012:BPM

[FN12a] Ana T. Freitas and Arcadi Navarro, editors. Bioinformatics for Personalized Medicine: 10th Spanish Symposium, JBI 2010, Torremolinos, Spain, October 27–29, 2010. Revised Selected Papers, volume 6620 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28061-7 (print), 3-642-28062-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28062-7.

Furia:2012:OMC

[FN12b] Carlo A. Furia and Sebastian Nanz, editors. Objects, Models, Components, Patterns: 50th International Conference, TOOLS 2012, Prague, Czech Republic, May 29– 31, 2012. Proceedings, volume 7304 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30560-1 (print), 3-642-30561-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30561-0.

Freksa:2008:SCV

[FNGW08] Christian Freksa, Nora S.

Newcombe, Peter Gärdenfors, and Stefan Wölfl, editors. Spatial Cognition VI. Learning, Reasoning, and Talking about Space: International Conference Spatial Cognition 2008, Freiburg, Germany, September 15– 19, 2008. Proceedings, volume 5248 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87600-6 (print), 3-540-87601-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87601-4.

Fox:2005:DLI

[FNPW05] Edward A. (Edward Alan)

Fox, Erich J. Neuhold, Pimrumpai Premsmit, and Vilas Wuwongse, editors. Digital libraries: implementing strategies and sharing experiences: 8th International Conference on Asian Digital Libraries, ICADL 2005, Bangkok, Thailand, December 12–15, 2005: proceedings, volume 3815 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30850-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4080 .I578 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30850-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3815.

Finin:1993:IKM

[FNY93] Timothy W. Finin, Charles K. Nicholas, and Yelena Yesha, editors. Information and knowledge management: expanding the definition of “database”: First International Conference, CIKM ’92, Baltimore, Maryland, USA, November 8–11, 1992: select papers, volume 752 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57419-0, 0-387-57419-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5586 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0752.
htm;   http://www.springerlink.
com/content/978-3-540-
57419-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=752.

Fisher:1995:EMT

[FO95] Michael Fisher and Richard Owens, editors. Executable modal and temporal logics: IJCAI ’93 Workshop, Chambery, France, August 28, 1993: proceedings, volume 897 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-58976-7 (Berlin), 0-387-58976-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I566 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0897.htm;
 http://www.springerlink.
com/content/978-3-540-
58976-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=897.

Updated papers presented at the Workshop on Executable Modal and Temporal Logics.

Fogarty:1994:ECA

[Fog94] Terence C. Fogarty, editor. Evolutionary computing: AISB Workshop, Leeds, UK, April 11– 13, 1994: Selected Papers, volume 865 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58483-8, 0-387-58483-

8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .A22 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0865.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58483-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=865.

Fogarty:1995:ECA

[Fog95] Terence C. Fogarty, editor. Evolutionary computing: AISB Workshop, Sheffield, UK, April 3– 4, 1995: Selected Papers, volume 993 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60469-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .A23 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0993.htm;
http://www.springerlink.
com/content/978-3-540-
60469-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=993.

Fogarty:1996:ECA

[Fog96] Terence C. Fogarty, editor. Evolutionary computing: AISB Workshop, Brighton, UK, April 1– 2, 1996: Selected Papers, volume 1143 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61749-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618.A23 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1143.
htm;   http://www.springerlink.
com/content/978-3-540-
61749-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1143.

Fierrez:2009:BIM

[FOGE^{+}09] Julian Fierrez, Javier Ortega-Garcia, Anna Esposito, Andrzej Drygajlo, and Marcos Faundez-Zanuy, editors. Biometric ID Management and Multimodal Communication: Joint COST 2101 and 2102 International Conference, BioID MultiComm 2009, Madrid, Spain, September 16–18, 2009. Proceedings, volume 5707 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04390-9 (print), 3-642-04391-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04391-8.

Fomin:2006:GTC

[Fom06] Fedor V. Fomin, editor. Graph-Theoretic Concepts in Computer Science: 32nd International Workshop, WG 2006, Bergen, Norway, June 22–24, 2006 Revised Papers, volume 4271 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48381-0 (print), 3-540-48382-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
48382-3.

Fonlupt:2000:AEE

[Fon00] Cyril Fonlupt, editor.

Artificial evolution: 4th European Conference, AE ’99, Dunkerque, France, November 3– 5, 1999: Selected Papers, volume 1829 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67846-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .A217 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1829.htm;
http://www.springerlink.
com/content/978-3-540-
67846-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1829.

Fonseca:2003:EMC

[Fon03] Carlos M. Fonseca, editor. Evolutionary multi-criterion optimization: Second International Conference, EMO 2003, Faro, Portugal, April 8–11, 2003: Proceedings, volume 2632 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-01869-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T57.95 .I57 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2632.
htm;   http://www.springerlink.
com/content/978-3-540-
01869-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2632.

Foo:1999:ATA

[Foo99] N. Y. Foo, editor. Advanced topics in artificial intelligence: 12th Australian Joint Conference on Artificial Intelligence, AI’99, Sydney, Australia, December 6–10, 1999: Proceedings, volume 1747 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66822-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1747. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1747.htm;
 http://www.springerlink.
com/content/978-3-540-
66822-0;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1747.

Ford:1988:SEE

[For88] Gary A. Ford, editor.

Software engineering education: proceedings, SEI conference 1988, Fairfax, Virginia, USA, April 28– 29, 1988, volume 327 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-96854-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.327. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0327.htm;
 http://www.springerlink.
com/content/978-0-387-
96854-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=327.

Proceedings of the annual SEI Conference on Software Engineering Education sponsored by the federally funded Software Engineering Institute (SEI) operated by Carnegie-Mellon University. Cf. pref. 1987 SEI conference published under the title: Issues in software engineering education.

Fossorier:1999:AAA

[Fos99] Marc Fossorier, editor. Applied algebra, algebraic algorithms, and errorcorrecting codes: 13th International Symposium, AAECC-13, Honolulu, Hawaii, USA, November 15–19, 1999: Proceedings, volume 1719 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66723-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1719; QA268 .A35 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1719.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66723-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1719.

Foster:2002:GPE

[Fos02] James A. Foster, editor. Genetic programming 5th European Conference, EuroGP 2002, Kinsale, Ireland, April 3–5, 2002: Proceedings, volume 2278 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43378-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.623 .E9 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2278.htm;
 http://www.springerlink.
com/content/978-3-540-
43378-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2278.

Floudas:1990:CTP

[FP90] Christodoulos A. Floudas and P. M. Pardalos. A collection of test problems for constrained global optimization algorithms, volume 455 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-53032-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 180 pp. LCCN TA331 .F58 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0455.htm;
 http://www.springerlink.
com/content/978-0-387-
53032-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=455.

Ferreira:1996:SCO

[FP96] Alfonso Ferreira and P. M. (Panos M.) Pardalos. Solving combinatorial optimization problems in parallel: methods and techniques, volume 1054 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61043-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). 274 pp. LCCN QA402.5 .S65 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1054.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61043-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1054.

Franch:2005:CBS

[FP05] Xavier Franch and Daniel Port, editors. COTSbased software systems: 4th international conference, ICCBSS 2005, Bilbao, Spain, February 7–11, 2005, proceedings, volume 3412 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24548-0 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .I33 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24548-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3412;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105900.

Forbrig:2008:EIS

[FP08] Peter Forbrig and Fabio Paternò, editors. Engineering Interactive Systems: Second Conference on Human-Centered Software Engineering, HCSE 2008, and 7th International Workshop on Task Models and Diagrams, TAMODIA 2008, Pisa, Italy, September 25–26, 2008. Proceedings, volume 5247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85991-8 (print), 3-540-85992-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85992-5.

Filler:2011:IHI

[FPCK11] Tomáš Filler, Tomáš Pevný, Scott Craver, and Andrew Ker, editors. Information Hiding: 13th International Conference, IH 2011, Prague, Czech Republic, May 18–20, 2011, Revised Selected Papers, volume 6958 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24177-8 (print), 3-642-24178-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24178-9.

Faltings:2005:RAC

[FPFR05] Boi Faltings, Adrian Petcu, François Fages, and Francesca Rossi, editors. Recent Advances in Constraints: Joint ERCIM/CoLogNET International Workshop on Constraint Solving and Constraint Logic Programming, CSCLP 2004, Lausanne, Switzerland, June 23–25, 2004: Revised Selected and Invited Papers, volume 3419 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25176-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .J65 2004. URL http:

//www.springerlink.
com/content/978-3-540-
25176-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3419;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106890.

Filliatre:2006:TPP

[FPMW06] Jean-Christophe Filliâtre, Christine Paulin-Mohring, and Benjamin Werner, editors. Types for proofs and programs: international workshop, TYPES 2004, Jouy-en-Josas, France, December 15– 18, 2004: revised selected papers, volume 3839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31428-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T96 2004.

URL http://www.springerlink.
com/content/978-3-540-
31428-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3839.

Fanelli:2011:FLA

[FPP11] Anna Maria Fanelli, Witold Pedrycz, and Alfredo Petrosino, editors. Fuzzy Logic and Applications: 9th International Workshop, WILF 2011, Trani, Italy, August 29–31,2011. Proceedings, volume 6857 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23712-6 (print), 3-642-23713-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23713-3.

Faratin:2004:AME

[FPRAW04] Peyman Faratin, David C. Parkes, Juan A. Rodràguez-Aguilar, and William E. Walsh, editors. Agent-Mediated Electronic Commerce V: Designing Mechanisms and Systems AAMAS 2003 workshop, AMEC 2003, Melbourne, Australia, July 15, 2003, Revised Selected Papers, volume 3048 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22674-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3048.htm;
 http://www.springerlink.
com/content/978-3-540-
22674-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3048;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99040.

Freund:2006:MCI

[FPRS06] Rudolf Freund, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa, editors. Membrane Computing: 6th International Workshop, WMC 2005, Vienna, Austria, July 18–21, 2005, Revised Selected and Invited Papers, volume 3850 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30948-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30948-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3850.

Feitelson:1995:JSS

[FR95a] Dror G. Feitelson and Larry Rudolph, editors. Job scheduling strategies for parallel processing: IPPS ’95 workshop, Santa Barbara, CA, USA, April 25, 1995: proceedings, volume 949 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60153-8 (Berlin: softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I59 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0949.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60153-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=949.

Ferreira:1995:PAI

[FR95b] Afonso Ferreira and Jose D. P. Rolim, editors. Parallel algorithms for irregularly structured problems: second International Workshop, IRREGULAR 95, Lyon, France, September 4–6, 1995: proceedings, volume 980 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60321-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .I59 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0980.htm;
 http://www.springerlink.
com/content/978-3-540-
60321-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=980.

Feitelson:1996:JSS

[FR96] Dror G. Feitelson and Larry Rudolph, editors. Job scheduling strategies for parallel processing: IPPS ’96 workshop, Honolulu, Hawaii, April 16, 1996: proceedings, volume 1162 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61864-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I59 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1162.
htm;   http://www.springerlink.
com/content/978-3-540-
61864-5;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1162.

Feitelson:1997:JSS

[FR97] Dror G. Feitelson and Larry Rudolph, editors. Job scheduling strategies for parallel processing: IPPS ’97 Workshop, Geneva, Switzerland, April 5, 1997: proceedings, volume 1291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63574-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I59 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1291.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63574-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1291.

Feitelson:1998:JSS

[FR98] Dror G. Feitelson and Larry Rudolph, editors. Job scheduling strategies for parallel processing: IPPS/SPDP ’98 Workshop, Orlando, Florida, USA, March 30, 1998: Proceedings, volume 1459 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64825-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I593 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1459.htm;
 http://www.springerlink.
com/content/978-3-540-
64825-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1459.

Flum:1999:CSL

[FR99a] Jorg Flum and M. (Mario)

Rodriguez Artalejo, editors. Computer science logic: 13th International Workshop, CSL ’99, 8th Annual Conference of the EACSL, Madrid, Spain, September 20–25, 1999: Proceedings, volume 1683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66536-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 W67 1999; QA75.5 .W624 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1683.
htm;   http://www.springerlink.
com/content/978-3-540-
66536-6;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1683.

France:1999:UML

[FR99b] Robert France and Bernhard Rumpe, editors. The unified modeling language: UML ’99: beyond the standard: Second International Workshop, Fort Collins, CO, October 28–30, 1999: Proceedings, volume 1723 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66712-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35 U54 1999; QA267.A1 L43 no.1723. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1723.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66712-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1723.

Feitelson:2000:JSS

[FR00] Dror G. Feitelson and Larry Rudolph, editors. Job Scheduling Strategies for Parallel Processing: IPDPS 2000 Workshop, JSSPP 2000, Cancun, Mexico, May 1, 2000. Proceedings, volume 1911 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41120-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1911. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1911.htm;
 http://www.springerlink.
com/content/978-3-540-
41120-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1911.

Papers presented at the sixth workshop on Job Scheduling Strategies for Parallel Processing, which was held in conjunction with the IPDPS 2000 Conference.

Feitelson:2001:JSS

[FR01a] Dror G. Feitelson and Larry Rudolph, editors. Job scheduling strategies for parallel processing: 7th International Workshop, JSSPP 2001, Cambridge, MA, USA, June 2001: Revised Papers, volume 2221 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42817-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.58 .J757 2001; QA76.58 .I59 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2221.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42817-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2221.

Ferreira:2001:SAS

[FR01b] Afonso Ferreira and Horst Reichel, editors. STACS 2001: 18th annual Symposium on Theoretical Aspects of Computer Science, Dresden, Germany, February 15–17, 2001: Proceedings, volume 2010 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41695-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S958 2001; QA267.A1 L43 no.2010. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2010.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41695-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2010.

Ferre:2009:FCA

[FR09] Sébastien Ferré and Sebastian Rudolph, editors. Formal Concept Analysis: 7th International Conference, ICFCA 2009 Darmstadt, Germany, May 21–24, 2009 Proceedings, volume 5548 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01814-9 (print), 3-642-01815-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01815-2.

Felber:2011:DAI

[FR11] Pascal Felber and Romain Rouvoy, editors. Distributed Applications and Interoperable Systems: 11th IFIP WG 6.1 International Conference, DAIS 2011, Reykjavik, Iceland, June 6– 9, 2011. Proceedings, volume 6723 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21386-3 (print), 3-642-21387-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21387-8.

Franz:1996:AAR

[Fra96] Alexander Franz. Automatic ambiguity resolution in natural language processing, volume 1171 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-62004-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 155 pp. LCCN QA76.9.N38F75 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1171.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62004-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1171.

Franklin:1999:FCT

[Fra99] Matthew Franklin, editor. Financial cryptography: Third International Conference, FC ’99, Anguilla, British West Indies, February 22–25, 1999: Proceedings, volume 1648 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66362-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1648.htm;
 http://www.springerlink.
com/content/978-3-540-
66362-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1648.

Frakes:2000:SRA

[Fra00] William B. Frakes, editor. Software reuse: advances in software reusability: 6th International Conference, ICSR-6, Vienna, Austria, June 27–29, 2000: Proceedings, volume 1844 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67696-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R47 I587 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1844.htm;
 http://www.springerlink.
com/content/978-3-540-
67696-6;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1844.

Frankel:2001:FCI

[Fra01] Yair Frankel, editor. Financial cryptography: 4th International Conference, FC 2000, Anguilla, British West Indies, February 20–24, 2000: Proceedings, volume 1962 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42700-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1962; HG1710 .F35 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1962.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42700-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1962.

Franklin:2004:ACC

[Fra04] Matt Franklin, editor. Advances in Cryptology — CRYPTO 2004: 24th Annual International Cryptology Conference, Santa Barbara, California, USA, August 15–19, 2004. Proceedings, volume 3152 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22668-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 2004.

URL http://www.springerlink.
com/content/978-3-540-
22668-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3152;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99099.

Faratin:2005:AME

[FRA05a] Peyman Faratin and Juan A. Rodríguez-Aguilar, editors. Agentmediated electronic commerce VI: theories for and engineering of distributed mechanisms and systems: AAMAS 2004 workshop, AMEC 2004, New York, NY, USA, July 19, 2004: revised selected papers, volume 3435 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29737-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        29737-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3435.

Fraigniaud:2005:DCI

[Fra05b] Pierre Fraigniaud, editor. Distributed computing: 19th international conference, DISC 2005, Cracow, Poland, September 26–29, 2005. proceedings, volume 3724 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29163-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29163-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3724.

Freytag:1987:TRQ

[Fre87] Johann Christoph Freytag. Translating relational queries into iterative programs, volume 261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-18000-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 131 pp. LCCN QA76.9.D3 F751 1987.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0261.
htm;   http://www.springerlink.
com/content/978-0-387-
18000-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=261.

Thesis (Ph. D.) — Harvard University, 1985.

Freuder:1996:PPC

[Fre96] Eugene C. Freuder, editor. Principles and practice of constraint programming—CP ’96: Second International Conference, CP ’96, Cambridge, MA, USA, August 19–22, 1996: proceedings, volume 1118 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61551-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .I58 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1118.htm;
 http://www.springerlink.
com/content/978-3-540-
61551-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1118.

Freitag:1998:TCL

[Fre98] Burkhard Freitag, editor. Transactions and change in logic databases: International Seminar on Logic Databases and the Meaning of Change, Schloss Dagstuhl, Germany, September 23– 27, 1996 and ILPS ’97 Post-Conference Workshop on (Trans)Actions and Change in Logic Programming and Deductive Databases, (DY-NAMICS ’97), Port Jefferson, NY, USA, October 17, 1997: invited surveys and Selected Papers, volume 1472 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65305-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5655 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1472.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65305-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1472.

Freivald:2001:FCT

[Fre01] R. V. Freivald, editor. Fundamentals of computation theory: 13th International Symposium, FCT 2001, Riga, Latvia, August 22–24, 2001: Proceedings, volume 2138 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42487-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2138; QA75.5 .F367 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2138.htm;
 http://www.springerlink.
com/content/978-3-540-
42487-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2138.

Freksa:2003:SCI

[Fre03] C. Freksa, editor. Spatial cognition III: routes and navigation, human memory and learning, spatial representation and spatial learning, volume 2685 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40430-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387 .S7 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2685.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40430-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2685.

Fritzson:1993:AAD

[Fri93] Peter A. Fritzson, editor. Automated and algorithmic debugging: first International Workshop, AADEBUG ’93, Linköping, Sweden, May 3–5, 1993: proceedings, volume 749 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57417-4 (Berlin), 0-387-57417-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D43 I54 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0749.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        57417-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=749.

Selection of papers presented at the First International Workshop on Automated and Algorithmic Debugging.

Fritzson:1994:CCI

[Fri94] Peter A. Fritzson, editor.

Compiler construction: 5th International Conference, CC ’94, Edinburgh, UK, April 7–9, 1994: proceedings, volume 786 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57877-3 (New York), 3-540-57877-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0786.htm;
http://www.springerlink.
com/content/978-0-387-
57877-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=786.

Fribourg:2001:CSL

[Fri01] Laurent Fribourg, editor.

Computer science logic: 15th International Workshop, CSL 2001, 10th Annual Conference of the EACSL, Paris, France, September 10–13, 2001: Proceedings, volume 2142 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42554-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W624 2001; QA267.A1 L43 no.2142. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2142.htm;
 http://www.springerlink.
com/content/978-3-540-
42554-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2142.

Fridrich:2004:IHI

[Fri04] Jessica Fridrich, editor. Information hiding: 6th international workshop, IH 2004, Toronto, Canada, May 23–25 2004, revised selected papers, volume 3200 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24207-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I49 2004. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3200;   http:
//www.springerlink.
com/content/978-3-540-
24207-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

0302-9743&volume=3200;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104759.

Frintrup:2006:VVA

[Fri06] Simone Frintrup, editor.

VOCUS: A Visual Attention System for Object Detection and Goal-Directed Search, volume 3899 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32759-2. ISSN 0302-9743 (print), 1611-3349 (electronic). 199 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32759-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3899.

Fonseca:2007:GSS

[FRL07] Frederico Fonseca, M. Andrea Rodríguez, and Sergei Levashkin, editors. GeoSpatial Semantics: Second International Conference, GeoS 2007, Mexico City, Mexico, November 29– 30, 2007. Proceedings, volume 4853 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76875-0 (print), 3-540-76876-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76876-0.

Frohner:2002:OOT

[Fro02] Akos Frohner, editor. Object-oriented technology: ECOOP 2001 workshop reader: ECOOP 2001 workshops, panels, and posters: Budapest, Hungary, June 18–22, 2001: Proceedings, volume 2323 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43675-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35 E33 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2323.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43675-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2323.

Feitelson:2002:JSS

[FRS02] Dror G. Feitelson, Larry Rudolph, and Uwe Schwiegelshohn, editors. Job scheduling strategies for parallel processing: 8th International Workshop, JSSPP 2002, Edinburgh, Scotland, UK, July 24, 2002: Revised Papers, volume 2537 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00172-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I59 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2537.htm;
 http://www.springerlink.
com/content/978-3-540-
00172-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2537.

Also available via the World Wide Web.

Feitelson:2003:JSS

[FRS03] Dror Feitelson, Larry Rudolph, and Uwe Schwiegelshohn, editors. Job Scheduling Strategies for Parallel Processing: 9th International Workshop, JSSPP 2003, Seattle, WA, USA, June 24, 2003: Revised Papers, volume 2862 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20405-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .J757 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2862.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20405-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2862;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b94203.

Feitelson:2005:JSSa

[FRS05] Dror G. Feitelson, Larry Rudolph, and Uwe Schwiegelshohn, editors. Job scheduling strategies for parallel processing: 10th international workshop, JSSPP 2004, New York, NY, USA, June 13, 2004: revised selected papers, volume 3277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25330-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .J757 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25330-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=

0302-9743&volume=3277;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107134.

Fages:2008:RAC

[FRS08] François Fages, Francesca Rossi, and Sylvain Soliman, editors. Recent Advances in Constraints: 12th Annual ERCIM International Workshop on Constraint Solving and Constraint Logic Programming, CSCLP 2007 Rocquencourt, France, June 7–8, 2007 Revised Selected Papers, volume 5129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89811-5 (print), 3-540-89812-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89812-2.

Frangi:2005:FIM

[FRSH05] Alejandro F. Frangi, Petia Ivanova Radeva, Andres Santos, and Monica Hernandez, editors. Functional Imaging and Modeling of the Heart: Third International Workshop, FIMH 2005, Barcelona, Spain, June 2–4, 2005. Proceedings, volume 3504 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26161-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP112.4 .I58 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26161-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3504.

Fabrikant:2010:GIS

[FRvKS10] Sara Irina Fabrikant, Tumasch Reichenbacher, Marc van Kreveld, and Christoph Schlieder, editors. Geographic Information Science: 6th International Conference, GIScience 2010, Zurich, Switzerland, September 14–17, 2010. Proceedings, volume 6292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15299-6 (print), 3-642-15300-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15300-6.

Fensel:1999:KAM

[FS99a] Dieter Fensel and Rudi Studer, editors. Knowledge acquisition modeling and management: 11th European Workshop, EKAW ’99, Dagstuhl Castle, Germany, May 26–29, 1999: Proceedings, volume 1621 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66044-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E92 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1621.htm;
http://www.springerlink.
com/content/978-3-540-
66044-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1621.

Fischer:1999:CLT

[FS99b] Paul Fischer and Hans Ulrich Simon, editors. Computational learning theory: 4th European Conference, Euro-COLT ’99, Nordkirchen, Germany, March 29– 31, 1999: Proceedings, volume 1572 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65701-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.7.E97 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1572.
htm;   http://www.springerlink.
com/content/978-3-540-
65701-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1572.

Fahringer:2003:ASA

[FS03] Thomas Fahringer and Bernhard Scholz, editors. Advanced symbolic analysis for compilers: new techniques and algorithms for symbolic program analysis and optimization, volume 2628 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-01185-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 .F34 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2628.htm;
 http://www.springerlink.
com/content/978-3-540-

01185-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2628.

Also available via the World Wide Web.

Fages:2005:PPS

[FS05] François Fages and Sylvain Soliman, editors. Principles and practice of Semantic Web reasoning: third international workshop, PPSWR 2005, Dagstuhl Castle, Germany, September 11– 16, 2005: proceedings, volume 3703 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28793-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815 .P67 2005.

URL http://www.springerlink.
com/content/978-3-540-
28793-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3703.

Furbach:2006:ART

[FS06] Ulrich Furbach and Natarajan Shankar, editors. Automated Reasoning: Third International Joint Conference, IJCAR 2006, Seattle, WA, USA, August 17–20, 2006. Proceedings, volume 4130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37187-7 (print), 3-540-37188-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37188-5.

Fasli:2007:AME

[FS07a] Maria Fasli and Onn Shehory, editors. Agent-Mediated Electronic Commerce. Automated Negotiation and Strategy Design for Electronic Markets: AAMAS 2006 Workshop, TADA/ AMEC 2006, Hakodate, Japan, May 9, 2006, Selected and Revised Papers, volume 4452 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72501-6 (print), 3-540-72502-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72502-2.

Fdida:2007:SIT

[FS07b] Serge Fdida and Kazunori Sugiura, editors. Sustainable Internet: Third Asian Internet Engineering Conference, AIN-TEC 2007, Phuket, Thailand, November 27– 29, 2007. Proceedings, volume 4866 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76808-4 (print), 3-540-76809-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76809-8.

Fiadeiro:2007:RTA

[FS07c] José Luiz Fiadeiro and Pierre-Yves Schobbens, editors. Recent Trends in Algebraic Development Techniques: 18th International Workshop, WADT 2006, La Roche en Ardenne, Belgium, June 1–3, 2006, Revised Selected Papers, volume 4409 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71997-0 (print), 3-540-71998-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        71998-4.

Frachtenberg:2007:JSS

[FS07d] Eitan Frachtenberg and Uwe Schwiegelshohn, editors. Job Scheduling Strategies for Parallel Processing: 12th International Workshop, JSSPP 2006, Saint-Malo, France, June 26, 2006, Revised Selected Papers, volume 4376 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71034-5 (print), 3-540-71035-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71035-6.

Frachtenberg:2008:JSS

[FS08] Eitan Frachtenberg and Uwe Schwiegelshohn, editors. Job Scheduling Strategies for Parallel Processing: 13th International Workshop, JSSPP 2007, Seattle, WA, USA, June 17, 2007. Revised Papers, volume 4942 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78698-8 (print), 3-540-78699-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78699-3.

Fiat:2009:AEA

[FS09a] Amos Fiat and Peter Sanders, editors. Algorithms, ESA 2009: 17th annual European symposium, Copenhagen, Denmark, September 7–9: proceedings, volume 5757 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CO-DEN LNCSD9. ISBN 3-642-04127-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 2009.

URL http://www.loc.
gov/catdir/enhancements/
fy1208/2009933188-d.
html;   http://www.
springerlink.com/content/
978-3-642-04127-3;
http://www.springerlink.
com/content/978-3-642-
04127-3/.

Frachtenberg:2009:JSS

[FS09b] Eitan Frachtenberg and Uwe Schwiegelshohn, editors. Job Scheduling Strategies for Parallel Processing: 14th International Workshop, JSSPP 2009, Rome, Italy, May 29, 2009. Revised Papers, volume 5798 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04632-0 (print), 3-642-04633-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04633-9.

Frachtenberg:2010:JSS

[FS10] Eitan Frachtenberg and Uwe Schwiegelshohn, editors. Job Scheduling Strategies for Parallel Processing: 15th International Workshop, JSSPP 2010, Atlanta, GA, USA, April 23, 2010, Revised Selected Papers, volume 6253 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16504-4 (print), 3-642-16505-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16505-4.

Falcidieno:2007:SMS

[FSA^{+}07] Bianca Falcidieno, Michela Spagnuolo, Yannis Avrithis, Ioannis Kompatsiaris, and Paul Buitelaar, editors. Semantic Multimedia: Second International Conference on Semantic and Digital Media Technologies, SAMT 2007, Genoa, Italy, December 5–7, 2007. Proceedings, volume 4816 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77033-X (print), 3-540-77051-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77051-0.

Ferrandez:2011:FNA

[FSdlPT11a] José Manuel Ferrández, José Ramón Álvarez Sánchez, Félix de la Paz, and F. Javier Toledo, editors. Foundations on Natural and Artificial Computation: 4th International Work-Conference on the Interplay Between Natural and Artificial Computation, IWINAC 2011, La Palma, Canary Islands, Spain, May 30 — June 3, 2011. Proceedings, Part I, volume 6686 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21343-X (print), 3-642-21344-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21344-1.

Ferrandez:2011:NCB

[FSdlPT11b] José Manuel Ferrández, José Ramón Álvarez Sánchez, Félix de la Paz, and F. Javier Toledo, editors. New Challenges on Bioinspired Applications: 4th International Workconference on the Interplay Between Natural and Artificial Computation, IWINAC 2011, La Palma, Canary Islands, Spain, May 30 — June 3, 2011. Proceedings, Part II, volume 6687 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21325-1 (print), 3-642-21326-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21326-7.

Fensel:2003:SWI

[FSM03] Dieter Fensel, Katia Sycara, and John Mylopoulos, editors. The Semantic Web—ISWC 2003: Second International Semantic Web Conference, Sanibel Island, FL, USA, October 20–23, 2003 proceedings, volume 2870 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20362-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2870.
htm;   http://www.springerlink.
com/content/978-3-540-
20362-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2870;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b14287.

Fishkin:2006:PCI

[FSN06] Kenneth P. Fishkin, Bernt Schiele, and Paddy Nixon, editors. Pervasive Computing: 4th International Conference, PER-VASIVE 2006, Dublin, Ireland, May 7–10, 2006: Proceedings, volume 3968 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33894-2 (paperback). LCCN QA76.59 .P47 2006. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33894-9.

Fritz:2009:CVS

[FSP09] Mario Fritz, Bernt Schiele, and Justus H. Piater, editors. Computer Vision Systems: 7th International Conference on Computer Vision Systems, ICVS 2009 Liège, Belgium, October 13– 15, 2009. Proceedings, volume 5815 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04666-5 (print), 3-642-04667-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04667-4.

Furnkranz:2006:KDD

[FSS06a]  Johannes  Fürnkranz,

Tobias Scheffer, and Myra Spiliopoulou, editors. Knowledge Discovery in Databases: PKDD 2006: 10th European Conference on Principles and Practice of Knowledge Discovery in Databases Berlin, Germany, September 18– 22, 2006 Proceedings, volume 4213 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45374-1 (print), 3-540-46048-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46048-0.

Furnkranz:2006:MLE

[FSS06b]  Johannes  Fürnkranz,

Tobias Scheffer, and Myra Spiliopoulou, editors. Machine Learning: ECML 2006: 17th European Conference on Machine Learning Berlin, Germany, September 18– 22, 2006 Proceedings, volume 4212 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45375-X (print), 3-540-46056-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46056-5.

Falcone:2001:TCS

[FST01] Rino Falcone, Munindar P. (Munindar Paul) Singh, and Yao-Hua Tan, editors. Trust in Cyber-societies: Integrating the Human and Artificial Perspectives, volume 2246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43069-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN BF575.T7 T78 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2246.htm;
 http://www.springerlink.
com/content/978-3-540-
43069-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2246.

Fisher:2009:CLM

[FST09] Michael Fisher, Fariba Sadri, and Michael Thielscher, editors. Computational Logic in Multi-Agent Systems: 9th International Workshop, CLIMA IX, Dresden, Germany, September 29– 30, 2008. Revised Selected and Invited Papers, volume 5405 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02733-4 (print), 3-642-02734-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02734-5.

Fratta:2009:NII

[FSTS09] Luigi Fratta, Henning Schulzrinne, Yutaka Takahashi, and Otto Spaniol, editors. NET-WORKING 2009: 8th International IFIP-TC 6 Networking Conference, Aachen, Germany, May 11–15, 2009. Proceedings, volume 5550 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01398-8 (print), 3-642-01399-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01399-7.

Foggia:2009:IAP

[FSV09] Pasquale Foggia, Carlo Sansone, and Mario Vento, editors. Image Analysis and Processing — ICIAP 2009: 15th International Conference Vietri sul Mare, Italy, September 8–11, 2009 Proceedings, volume 5716 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04145-0 (print), 3-642-04146-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04146-4.

Fribourg:1994:LPS

[FT94] Laurent Fribourg and Franco Turini, editors. Logic program synthesis and transformation

— Meta-programming in Logic: 4th International Workshops, LOPSTR ’94 and META ’94, Pisa, Italy, June 20–20, 1994: proceedings, volume 883 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58792-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73 .I57 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0883.htm;
 http://www.springerlink.
com/content/978-3-540-
58792-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=883.

Fleischer:2004:ACI

[FT04a] Rudolf Fleischer and Gerhard Trippen, editors. Algorithms and computation: 15th international symposium, ISAAC 2004, Hong Kong, China, December 20–22, 2004: Proceedings, volume 3341 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24131-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I74 2004. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3341;   http:
//www.springerlink.
com/content/978-3-540-
24131-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3341;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104582.

Frederking:2004:MTR

[FT04b] Robert E. Frederking and Kathryn B. Taylor, editors. Machine Translation: From Real Users to Research: 6th Conference of the Association for Machine Translation in the Americas, AMTA 2004: Washington, DC, USA, September 28–October 2, 2004: Proceedings, volume 3265 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23300-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3265.htm;
 http://www.springerlink.
com/content/978-3-540-
23300-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3265;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100780.

Fahrenberg:2011:FMA

[FT11] Uli Fahrenberg and Stavros Tripakis, editors. Formal Modeling and Analysis of Timed Systems: 9th International Conference, FORMATS 2011, Aalborg, Denmark, September 21–23, 2011. Proceedings, volume 6919 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24309-6 (print), 3-642-24310-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24310-3.

Fischer:2006:MST

[FTAZ06] Klaus Fischer, Ingo J.

Timm, Elisabeth André, and Ning Zhong, editors. Multiagent System Technologies: 4th German Conference, MATES 2006 Erfurt, Germany, September 19–20, 2006 Proceedings, volume 4196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45376-8 (print), 3-540-46057-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46057-2.

Fyfe:2010:IDE

[FTC^{+}10] Colin Fyfe, Peter Tino, Darryl Charles, Cesar Garcia-Osorio, and Hujun Yin, editors. Intelligent Data Engineering and Automated Learning — IDEAL 2010: 11th International Conference, Paisley, UK, September 1–3, 2010. Proceedings, volume 6283 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15380-1 (print), 3-642-15381-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15381-5.

Furukawa:1988:LPP

[FTF88] Koichi Furukawa, H. Tanaka, and T. Fujisaki, editors. Logic programming ’87: proceedings of the 6th conference, Tokyo, Japan, June 22– 24, 1987, volume 315 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19426-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L587621 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0315.htm;
 http://www.springerlink.
com/content/978-0-387-
19426-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=315.

“Papers presented at the Sixth Logic Programming Conference”–Fwd.

Furukawa:1989:LPP

[FTF89] Koichi Furukawa, H. Tanaka, and T. Fujisaki, editors. Logic programming ’88: proceedings of the 7th conference, Tokyo, Japan, April 11– 14, 1988, volume 383 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51564-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .L63 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0383.htm;
http://www.springerlink.
com/content/978-0-387-
51564-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=383.

“Papers presented at the Seventh Logic Programming Conference”– Foreword.

Furukawa:1991:LPP

[FTF91] Koichi Furukawa, H. Tanaka, and T. Fujisaki, editors. Logic Programming ’89: proceedings of the 8th Conference, Tokyo, Japan, July 12– 14, 1989, volume 485 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-53919-0, 0-387-53919-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 L63 1991. DM18.86.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0485.
htm;   http://www.springerlink.
com/content/978-3-540-
53919-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=485.

Forsyth:2008:CVEa

[FTZ08a] David Forsyth, Philip Torr, and Andrew Zisserman, editors. Computer Vision — ECCV 2008: 10th European Conference on Computer Vision, Marseille, France, October 12–18, 2008, Proceedings, Part I, volume 5302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88681-8 (print), 3-540-88682-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
88682-2.

Forsyth:2008:CVEb

[FTZ08b] David Forsyth, Philip Torr, and Andrew Zisserman, editors. Computer Vision — ECCV 2008: 10th European Conference on Computer Vision, Marseille, France, October 12–18, 2008, Proceedings, Part II, volume 5303 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88685-0 (print), 3-540-88688-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88688-4.

Forsyth:2008:CVEc

[FTZ08c] David Forsyth, Philip Torr, and Andrew Zisserman, editors. Computer Vision — ECCV 2008: 10th European Conference on Computer Vision, Marseille, France, October 12–18, 2008, Proceedings, Part III, volume 5304 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88689-3 (print), 3-540-88690-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88690-7.

Forsyth:2008:CVEd

[FTZ08d] David Forsyth, Philip Torr, and Andrew Zisserman, editors. Computer Vision — ECCV 2008: 10th European Conference on Computer Vision, Marseille, France, October 12–18, 2008, Proceedings, Part IV, volume 5305 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88692-3 (print), 3-540-88693-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88693-8.

Furuhashi:1996:FLN

[FU96] T. (Takeshi) Furuhashi and Y. (Yoshiki) Uchikawa, editors. Fuzzy logic, neural networks, and evolutionary computation: IEEE/Nagoya University World Wisepersons Workshop, Nagoya, Japan, November 14– 15, 1995: Selected Papers, volume 1152 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61988-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87.I346 1995. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1152.
htm;   http://www.springerlink.
com/content/978-3-540-
61988-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1152.

Fuchs:1998:LPS

[Fuc98] N. E. Fuchs, editor.

Logic program synthesis and transformation: 7th International Workshop, LOPSTR’97, Leuven, Belgium, July 10– 12, 1997: Proceedings, volume 1463 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65074-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1463. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1463.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65074-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1463.

Fuchs:2010:CNL

[Fuc10] Norbert E. Fuchs, editor. Controlled Natural Language: Workshop on Controlled Natural Language, CNL 2009, Marettimo Island, Italy, June 8–10, 2009. Revised Papers, volume 5972 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14417-9 (print), 3-642-14418-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14418-9.

Fuller:1975:ADD

[Ful75] Samuel H. Fuller. Analysis of drum and disk storage units, volume 31 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CO-DEN LNCSD9. ISBN 0-387-07186-5. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 283 pp. LCCN TK7895.M3 F841. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0031.htm;
 http://www.springerlink.
com/content/978-0-387-
07186-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=31.

Fumy:1997:ACE

[Fum97] Walter Fumy, editor.

Advances in cryptology

— EUROCRYPT ’97: International Conference on the Theory and Application of Cryptographic Techniques, Konstanz, Germany, May 11–15, 1997: proceedings, volume 1233 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62975-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1233.
htm;   http://www.springerlink.
com/content/978-3-540-
62975-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1233.

Sponsored by the International Association for Cryptologic Research (IACR).

Furuhashi:1995:AFL

[Fur95] T. Furuhashi, editor. Advances in fuzzy logic, neural networks, and genetic algorithms: IEEE/ Nagoya-University World Wisepersons Workshop, Nagoya, Japan, August 9–10, 1994: Selected Papers, volume 1011 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60607-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I346 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1011.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60607-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1011.

Furbach:2005:KAA

[Fur05] Ulrich Furbach, editor. KI 2005: advances in artificial intelligence: 28th Annual German Conference on AI, KI 2005, Koblenz, Germany, September 11– 14, 2005: proceedings, volume 3698 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28761-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G466 2005; Q334 .G47 2005. URL http:

//www.springerlink.
com/content/978-3-540-
28761-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3698.

Falsafi:2001:PAC

[FV01] Babak Falsafi and T. N.

Vijaykumar, editors. Power-aware computer systems: First International Workshop, PACS 2000, Cambridge, MA, USA, November 12, 2000: Revised Papers, volume 2008 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42329-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.P68 I58 2000; QA267.A1 L43 no.2008. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2008.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        42329-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2008.

Falsafi:2003:PAC

[FV03] Babak Falsafi and T. N. Vijaykumar, editors. Power-aware computer systems: Second International Workshop, PACS 2002, Cambridge, MA, USA, February 2, 2002: Revised Papers, volume 2325 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-01028-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.P68 I58 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2325.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        01028-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2325.

Falsafi:2004:PAC

[FV04] Babak Falsafi and T. N. Vijaykumar, editors. Power-aware computer systems: Third International Workshop, PACS 2003, San Diego, CA, USA, December 1, 2003: revised papers, volume 3164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-24031-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.P68 P33 2003. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3164;   http:
//www.springerlink.
com/content/978-3-540-
24031-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3164;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104218.

Falsafi:2005:PAC

[FV05] Babak Falsafi and T. N.

Vijaykumar, editors. Power-aware computer systems: 4th International Workshop, PACS 2004, Portland, OR, USA, December 5, 2004: revised selected papers, volume 3471 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29790-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.P68 .P33 2004; TK7895.P68 .P33. URL

                                        http://www.springerlink.
                                        com/content/978-3-540-
                                        29790-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3471.

Field:2009:CML

[FV09] John Field and Vasco T. Vasconcelos, editors. Coordination Models and Languages: 11th International Conference, COORDINATION 2009, Lisboa, Portugal, June 9–12, 2009. Proceedings, volume 5521 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02052-6 (print), 3-642-02053-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02053-7.

Fermuller:2010:LPA

[FV10] Christian G. Fermüller and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 17th International Conference, LPAR-17, Yogyakarta, Indonesia, October 10–15, 2010. Proceedings, volume 6397 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16241-X (print), 3-642-16242-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16242-8.

Fukuda:2010:MSI

[FvdHJT10] Komei Fukuda, Joris van der Hoeven, Michael Joswig, and Nobuki Takayama, editors. Mathematical software — ICMS 2010: third International Congress on Mathematical Software, Kōbe, Japan, September 13–17, 2010: proceedings, volume 6327 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-15581-2 (paperback), 3-642-15582-0 (ebook). LCCN QA76.95 .I5654 2010. URL http:

//www.springerlink.
com/content/978-3-642-
15581-9.

Fisher:2006:LAI

[FvdHKL06] Michael Fisher, Wiebe van der Hoek, Boris Konev, and Alexei Lisitsa, editors. Logics in Artificial Intelligence: 10th European Conference, JELIA 2006 Liverpool, UK, September 13–15, 2006 Proceedings, volume 4160 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39625-X (print), 3-540-39627-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
39627-7.

Fisher:2012:CLM

[FvdTDG12] Michael Fisher, Leon van der Torre, Mehdi Dastani, and Guido Governatori, editors. Computational Logic in Multi-Agent Systems: 13th International Workshop, CLIMA XIII, Montpellier, France, August 27– 28, 2012. Proceedings, volume 7486 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32896-2 (print), 3-642-32897-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32897-8.

Fronhofer:1992:PIS

[FW92] B. Fronhofer and Graham Wrightson, editors. Parallelization in inference systems: international workshop, Dagstuhl Castle, Germany, December 17– 18, 1990: proceedings, volume 590 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55425-4 (Berlin), 0-387-55425-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P379 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0590.htm;
 http://www.springerlink.
com/content/978-3-540-
55425-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=590.

Fiat:1998:OAS

[FW98] Amos Fiat and Gerhard Woeginger. Online algorithms: the state of the art, volume 1442 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64917-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 436 pp. LCCN QA76.9.A43 F53 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1442.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64917-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1442.

Fischetti:2007:IPC

[FW07] Matteo Fischetti and David P. Williamson, editors. Integer Programming and Combinatorial Optimization: 12th International IPCO Conference, Ithaca, NY, USA, June 25–27, 2007. Proceedings, volume 4513 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72791-4 (print), 3-540-72792-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72792-7.

Franceschinis:2009:ATP

[FW09] Giuliana Franceschinis and Karsten Wolf, editors. Applications and Theory of Petri Nets: 30th International Conference, PETRI NETS 2009, Paris, France, June 22–26, 2009. Proceedings, volume 5606 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02423-8 (print), 3-642-02424-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02424-5.

Fan:2005:AWA

[FWY05] Wenfei Fan, Zhaohui Wu, and Jun Yang, editors. Advances in Web-Age Information Management: 6th International Conference, WAIM 2005, Hangzhou, China, October 11–13, 2005. Proceedings, volume 3739 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29227-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29227-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3739.

Feng:2006:WIS

[FWZH06] Ling Feng, Guoren Wang, Cheng Zeng, and Ruhua Huang, editors. Web Information Systems — WISE 2006 Workshops: WISE 2006 International Workshops, Wuhan, China, October 23–26, 2006. Proceedings, volume 4256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47663-6 (print), 3-540-47664-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        47664-1.

Fleischer:2008:AAI

[FX08] Rudolf Fleischer and Jinhui Xu, editors. Algorithmic Aspects in Information and Management: 4th International Conference, AAIM 2008, Shanghai, China, June 23–25, 2008. Proceedings, volume 5034 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68865-X (print), 3-540-68880-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68880-8.

Foresti:2012:CSE

[FYM12] Sara Foresti, Moti Yung, and Fabio Martinelli, editors. Computer Security — ESORICS 2012: 17th European Symposium on Research in Computer Security, Pisa, Italy, September 10–12, 2012. Proceedings, volume 7459 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33166-1 (print), 3-642-33167-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33167-1.

Franchi-Zannettacci:1986:CCT

[FZ86] P. Franchi-Zannettacci, editor. CAAP 86: 11th Colloquium on Trees in Algebra and Programming, Nice, France, March 24–26, 1986: proceedings, volume 214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16443-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166.2 .C651 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0214.htm;
 http://www.springerlink.
com/content/978-0-387-
16443-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=214.

Figueiredo:2001:EMM

[FZJ01] Mario Figueiredo, Josiane Zerubia, and Anil K. Jain, editors. Energy minimization methods in computer vision and pattern recognition: Third International Workshop, EMM-CVPR 2001, Sophia Antipolis, France, September 3–5, 2001: Proceedings, volume 2134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42523-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E54 2001; QA267.A1 L43 no.2134. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2134.htm;
 http://www.springerlink.
com/content/978-3-540-

42523-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2134.

Faundez-Zanuy:2005:NAA

[FZJE^{+}05] Marcos Faundez-Zanuy, Léonard Janer, Anna Esposito, Antonio Satue-Villar, Josep Roure, and Virginia Espinosa-Duro, editors. Nonlinear Analyses and Algorithms for Speech Processing: International Conference on Non-Linear Speech Processing, NOLISP 2005, Barcelona, Spain, April 19–22, 2005, Revised Selected Papers, volume 3817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-31257-9 (print), 3-540-32586-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32586-4.

Faundez-Zanuy:2006:NAA

[FZJE^{+}06] Marcos Faundez-Zanuy, Léonard Janer, Anna Esposito, Antonio Satue-Villar, Josep Roure, and Virginia Espinosa-Duro, editors. Nonlinear Analyses and Algorithms for Speech Processing: International Conference on Non-Linear Speech Processing, NOLISP 2005, Barcelona, Spain, April 19–22, 2005, Revised Selected Papers, volume 3817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31257-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31257-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3817.

Goldberg:2009:PET

[GA09] Ian Goldberg and Mikhail J. Atallah, editors. Privacy Enhancing Technologies: 9th International Symposium, PETS 2009, Seattle, WA, USA, August 5–7, 2009. Proceedings, volume 5672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03167-6 (print), 3-642-03168-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
03168-7.

Gibbons:2006:DCS

[GAA06] Phil Gibbons, Tarek Abdelzaher, and James Aspnes, editors. Distributed Computing in Sensor Systems: Second IEEE International Conference, DCOSS 2006, San Francisco, CA, USA, June 18–20, 2006. Proceedings, volume 4026 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35227-9 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35227-3.

Gabbay:1997:QQP

[Gab97] Dov M. Gabbay, editor.

Qualitative and quantitative practical reasoning: First International Joint Conference on Qualitative and Quantitative Practical Reasoning ECSQARU-FAPR’97, Bad Honnef, Germany, June 9– 12, 1997, proceedings, volume 1244 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63095-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q339.25 .I58 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1244.htm;
 http://www.springerlink.
com/content/978-3-540-
63095-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1244.

Gregori:2002:ALN

[GAB02] Enrico Gregori, Giuseppe Anastasi, and Stefano Basagni, editors. Advanced Lectures on Networking: NETWORK-ING 2002 Tutorials, volume 2497 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00165-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .A327 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2497.htm;
 http://www.springerlink.
com/content/978-3-540-
00165-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2497.

Rev. version of 7 tutorials given at the Networking 2002 Conference, Pisa, Italy, May 2002.

Gaborit:2013:PQC

[Gab13] Philippe Gaborit, editor.

Post-quantum cryptography: 5th International Workshop, PQCRYPTO 2013, Limoges, France, June 4–7, 2013. Proceedings, volume 7932 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2013. ISBN 3-642-38615-6 (paperback), 3-642-38616-4 (e-book). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
38616-9.

Gacek:2002:SRM

[Gac02] Cristina Gacek, editor.

Software reuse: methods, techniques, and tools: 7th International Conference, ICSR-7, Austin, TX, USA, April 15– 19, 2002: Proceedings, volume 2319 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43483-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R47 I59 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2319.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43483-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2319.

Gaede:1996:CDA

[Gae96] V. Gaede, editor. Constraint databases and applications: Second International Workshop on Constraint Database and Systems, CDB ’97, Delphi, Greece, January 11– 12, 1997, CP ’96 Workshop on Constraints and Databases, Cambridge, MA, USA, August 19, 1996: Selected Papers, volume 1191 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-62501-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .I59 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1191.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62501-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1191.

Gallagher:1997:LPS

[Gal97a] John (John P.) Gallagher, editor. Logic program synthesis and transformation: 6th International Workshop, LOPSTR’96, Stockholm, Sweden, August 28– 30, 1996: proceedings, volume 1207 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62718-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I57 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1207.htm;
 http://www.springerlink.
com/content/978-3-540-
62718-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1207.

Galmiche:1997:ARA

[Gal97b] Didier Galmiche, editor. Automated reasoning with analytic tableaux and related methods: International Conference, TABLEAUX ’97, Ponta-Mousson, France, May 13–16, 1997: proceedings, volume 1227 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62920-3 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T33 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1227.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62920-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1227.

Galbraith:2007:CCI

[Gal07] Steven D. Galbraith, editor. Cryptography and Coding: 11th IMA International Conference, Cirencester, UK, December 18–20, 2007. Proceedings, volume 4887 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77271-5 (print), 3-540-77272-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77272-9.

Garcia-Alfaro:2012:FPS

[GAL12] Joaquin Garcia-Alfaro and Pascal Lafourcade, editors. Foundations and Practice of Security: 4th Canada-France MI-TACS Workshop, FPS 2011, Paris, France, May 12–13, 2011, Revised Selected Papers, volume 6888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27900-7 (print), 3-642-27901-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27901-0.

Gross:2011:TAR

[GAM^{+}11] Roderich Groß, Lyuba Alboul, Chris Melhuish, Mark Witkowski, Tony J. Prescott, and Jacques Penders, editors. Towards Autonomous Robotic Systems: 12th Annual Conference, TAROS 2011, Sheffield, UK, August 31 – September 2, 2011. Proceedings, volume 6856 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23231-0 (print), 3-642-23232-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23232-9.

Ganzinger:1988:EES

[Gan88] H. Ganzinger, editor. ESOP ’88: 2nd European Symposium on Programming, Nancy, France, March 21–24, 1988: proceedings, volume 300 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19027-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E9761 1988. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0300.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        19027-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=300.

Held in conjunction with

CAAP ’88.

Ganzinger:1996:RTA

[Gan96] H. (Harald) Ganzinger, editor. Rewriting techniques and applications: 7th international conference, RTA-96, New Brunswick, NJ, USA, July 27–30, 1996: proceedings, volume 1103 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61464-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .R48 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1103.htm;
 http://www.springerlink.
com/content/978-3-540-
61464-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1103.

Ganzinger:1999:ADC

[Gan99] H. Ganzinger, editor.

Automated deduction — CADE 16: 16th International Conference on Automated Deduction, Trento, Italy, July 7–10, 1999: Proceedings, volume 1632 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66222-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I57 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1632.htm;
 http://www.springerlink.
com/content/978-3-540-
66222-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1632.

Gschwind:2005:SCI

[GAN05] Thomas Gschwind, Uwe Aßmann, and Oscar Nierstrasz, editors. Software composition: 4th international workshop, SC 2005, Edinburgh, UK, April 9, 2005. revised selected papers, volume 3628 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28748-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28748-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3628.

Garcia-Alfaro:2012:DPM

[GANACBdCdV12] Joaquin Garcia-Alfaro, Guillermo Navarro-Arribas, Nora Cuppens-Boulahia, and Sabrina de Capitani di Vimercati, editors. Data Privacy Management and Autonomous Spontaneous Security: 6th International Workshop, DPM 2011, and 4th International Workshop, SE-TOP 2011, Leuven, Belgium, September 15–16, 2011, Revised Selected Papers, volume 7122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28878-2 (print), 3-642-28879-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28879-1.

Garcia-Alfaro:2010:DPM

[GANACBR10] Joaquin Garcia-Alfaro, Guillermo Navarro-Arribas, Nora Cuppens-Boulahia, and Yves Roudier, editors. Data Privacy Management and Autonomous Spontaneous Security: 4th International Workshop, DPM 2009 and Second International Workshop, SE-TOP 2009, St. Malo, France, September 24– 25, 2009, Revised Selected Papers, volume 5939 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11206-4 (print), 3-642-11207-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        11207-2.

Garcia-Alfaro:2011:DPM

[GANACL11] Joaquin Garcia-Alfaro, Guillermo Navarro-Arribas, Ana Cavalli, and Jean Leneutre, editors. Data Privacy Management and Autonomous Spontaneous Security: 5th International Workshop, DPM 2010 and 3rd International Workshop, SETOP 2010, Athens, Greece, September 23, 2010, Revised Selected Papers, volume 6514 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19347-1 (print), 3-642-19348-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19348-4.

Guelfi:2003:SED

[GAR03] Nicolas Guelfi, E. Astesiano, and Gianna Reggio, editors. Scientific engineering for distributed Java applications: International Workshop, FIDJI 2002. Luxembourg-Kirchberg, Luxembourg, November 28–29, 2002: Revised Papers, volume 2604 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00679-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F53 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2604.htm;
 http://www.springerlink.
com/content/978-3-540-
00679-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2604.

Guelfi:2004:SED

[GAR04] Nicolas Guelfi, Egidio Astesiano, and Gianna Reggio, editors. Scientific Engineering of Distributed Java Applications: Third International Workshop, FIDJI 2003, Luxembourg-Kirchberg, Luxembourg, November 27–28, 2003: Revised Papers, volume 2952 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21091-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F53 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2952.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21091-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2952;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95352.

Gauthier:1993:AEA

[Gau93] Michel J. Gauthier, editor. Ada-Europe ’93: 12th Ada-Europe International Conference, “Ada Sans Frontieres”, Paris, France, June 14–18, 1993: proceedings, volume 688 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56802-6 (Berlin), 0-387-56802-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 A24 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0688.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        56802-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=688.

Goto:1986:RSS

[GAY86] Eiichi Goto, Keijiro Araki, and T. Yuasa, editors. RIMS symposia on software science and engineering II: proceedings of the symposia 1983 and 1984, Kyoto, Japan, volume 220 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16470-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .R541 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0220.htm;
 http://www.springerlink.
com/content/978-0-387-
16470-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=220.

Gerbier:1977:MPC

[GB77] A. Gerbier and P. Bacchus. Mes premières constructions de programmes. (French) [My First Programs], volume 55 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CODEN LNCSD9. ISBN 3-540-08438-X. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 256 pp. LCCN QA76.6 .G47. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t0055.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        08438-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=55.

Geppert:1997:RDS

[GB97] Andreas Geppert and Mikael Berndtsson, editors. Rules in database systems: Third International Workshop, RIDS ’97 Skövde, Sweden, June 26–28, 1997, proceedings, volume 1312 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63516-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3R854 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1312.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63516-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1312.

Garijo:1999:MAS

[GB99] Francisco J. Garijo and Magnus Boman, editors. Multi-agent system engineering: 9th European Workshop on Modelling Autonomous Agents in a Multi-Agent World, MAAMAW’99: Valencia, Spain, June 30– July 2, 1999: Proceedings, volume 1647 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66281-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1647. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1647.htm;
 http://www.springerlink.
com/content/978-3-540-
66281-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1647.

Guelfi:2007:RIS

[GB07] Nicolas Guelfi and Didier Buchs, editors. Rapid Integration of Software Engineering Techniques: Third International Workshop, RISE 2006, Geneva, Switzerland, September 13–15, 2006. Revised Selected Papers, volume 4401 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71875-3 (print), 3-540-71876-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71876-5.

Giacobini:2008:AEC

[GBC^{+}08] Mario Giacobini, Anthony Brabazon, Stefano Cagnoni, Gianni A. Di Caro, Rolf Drechsler, Anikó Ekárt, Anna Isabel Esparcia-Alcázar, Muddassar Farooq, Andreas Fink, Jon McCormack, Michael O’Neill, Juan Romero, Franz Rothlauf, Giovanni Squillero, A. Şima Uyar, and Shengxiang Yang, editors. Applications of Evolutionary Computing: EvoWorkshops 2008: EvoCOMNET, EvoFIN, EvoHOT, EvoIASP, EvoMUSART, EvoNUM, EvoSTOC, and Evo-TransLog, Naples, Italy, March 26–28, 2008. Proceedings, volume 4974 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78760-7 (print), 3-540-78761-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
78761-7.

Giacobini:2009:AEC

[GBC^{+}09] Mario Giacobini, Anthony Brabazon, Stefano Cagnoni, Gianni A. Di Caro, Anikó Ekárt, Anna Isabel Esparcia-Alcázar, Muddassar Farooq, Andreas Fink, and Penousal Machado, editors. Applications of Evolutionary Computing: EvoWorkshops 2009: EvoCOMNET, EvoENVIRONMENT, EvoFIN, EvoGAMES, EvoHOT, EvoIASP, EvoIN-TERACTION, EvoMUSART, EvoNUM, EvoSTOC, EvoTRANSLOG, Tübingen, Germany, April 15– 17, 2009. Proceedings, volume 5484 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01128-4 (print), 3-642-01129-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01129-0.

Garbow:1977:MER

[GBDM77] B. S. Garbow, J. M.

Boyle, J. J. Dongarra, and C. B. Moler. Matrix eigensystem routines: EISPACK guide extension, volume 51 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CO-DEN LNCSD9. ISBN 0-387-08254-9. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 343 pp. LCCN QA193 .M381. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0051.htm;
 http://www.springerlink.
com/content/978-0-387-
08254-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=51.

Supplements the second edition of Matrix eigensystem routines: EIS-PACK guide [SBD^{+}76].

Gama:2011:AID

[GBH11] João Gama, Elizabeth Bradley, and Jaakko Hollmén, editors. Advances in Intelligent Data Analysis X: 10th International Symposium, IDA 2011, Porto, Portugal, October 29– 31, 2011. Proceedings, volume 7014 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24799-7 (print), 3-642-24800-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24800-9.

Gradmann:2011:RAT

[GBMS11] Stefan Gradmann, Francesca Borri, Carlo Meghini, and Heiko Schuldt, editors. Research and Advanced Technology for Digital Libraries: International Conference on Theory and Practice of Digital Libraries, TPDL 2011, Berlin, Germany, September 26–28, 2011. Proceedings, volume 6966 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24468-8 (print), 3-642-24469-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24469-8.

Gruska:1981:MFC

[GC81] Jozef Gruska and Michal Chytil, editors. Mathematical foundations of computer science 1981: proceedings, 10th symposium, Strbske Pleso, Czechoslovakia, August 31–September 4, 1981, volume 118 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10856-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .M38.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0118.
htm;   http://www.springerlink.
com/content/978-0-387-
10856-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=118.

Papers presented at the 10th Symposium on Mathematical Foundations of Computer Science.

Gama:2005:MLE

[GCB^{+}05] João Gama, Rui Camacho, Pavel B. Brazdil, Alípio M. Jorge, and Luís Torgo, editors. Machine Learning: ECML 2005: 16th European Conference on Machine Learning, Porto, Portugal, October 3–7, 2005. Proceedings, volume 3720 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29243-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29243-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3720.

Gregori:2002:WEP

[GCC^{+}02a] Enrico Gregori, Ludmila Cherkasova, Gianpaolo Cugola, Fabio Panzieri, and Gian P. Picco, editors. Web Engineering and Peer-to-Peer Computing: NETWORK-ING 2002 Workshops, Pisa, Italy, May 19– 24, 2002. Revised Papers, volume 2376 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44177-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .N4814 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2376.htm;
http://www.springerlink.
com/content/978-3-540-
44177-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2376.

Gregori:2002:NNT

[GCC^{+}02b] Enrico Gregori, Marco Conti, Andrew T. Campbell, Guy Omidyar, and Moshe Zukerman, editors. NETWORKING 2002. Networking Technologies, Services, and Protocols; Performance of Computer and Communication Networks; Mobile and Wireless Communications: Second International IFIP-TC6 Networking Conference, Pisa, Italy, May 19–24, 2002. Proceedings, volume 2345 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43709-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .N49 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2345.htm;
 http://www.springerlink.
com/content/978-3-540-
43709-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2345.

Garcia-Castro:2012:SWE

[GCFA12] Raúl García-Castro, Dieter Fensel, and Grigoris Antoniou, editors. The Semantic Web: ESWC 2011 Workshops: ESWC 2011 Workshops, Heraklion, Greece, May 29–30, 2011, Revised Selected Papers, volume 7117 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25952-9 (print), 3-642-25953-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25953-1.

Gama:2009:DSI

[GCJB09] João Gama, Vítor Santos Costa, Alípio Mário Jorge, and Pavel B. Brazdil, editors. Discovery Science: 12th International Conference, DS 2009, Porto, Portugal, October 3–5, 2009, volume 5808 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04746-7 (print), 3-642-04747-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04747-3.

Gibet:2006:GHC

[GCK06] Sylvie Gibet, Nicolas Courty, and Jean-François Kamp, editors. Gesture in Human-Computer Interaction and Simulation: 6th International Gesture Workshop, GW 2005, Berder Island, France, May 18–20, 2005, Revised Selected Papers, volume 3881 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32624-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32624-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3881.

Garcia:2006:SEM

[GCL^{+}06] Alessandro Garcia, Ricardo Choren, Carlos Lucena, Paolo Giorgini, Tom Holvoet, and Alexander Romanovsky, editors. Software Engineering for Multi-Agent Systems IV: Research Issues and Practical Applications, volume 3914 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33580-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 255 pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33580-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3914.

Goodall:2008:VCS

[GCM08] John R. Goodall, Gregory Conti, and Kwan-Liu Ma, editors. Visualization for Computer Security: 5th International Workshop, VizSec 2008, Cambridge, MA, USA, September 15, 2008. Proceedings, volume 5210 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85931-4 (print), 3-540-85933-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85933-8.

Goh:2007:ADL

[GCSR07] Dion Hoe-Lian Goh, Tru Hoang Cao, Ingeborg Torvik Sølvberg, and Edie Rasmussen, editors. Asian Digital Libraries. Looking Back 10 Years and Forging New Frontiers: 10th International Conference on Asian Digital Libraries, ICADL 2007, Hanoi, Vietnam, December 10–13, 2007. Proceedings, volume 4822 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77093-3 (print), 3-540-77094-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77094-7.

Garay:2010:SCN

[GD10] Juan A. Garay and Roberto De Prisco, editors. Security and Cryptography for Networks: 7th International Conference, SCN 2010, Amalfi, Italy, September 13–15, 2010. Proceedings, volume 6280 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15316-X (print), 3-642-15317-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15317-4.

Gelbukh:2005:MAA

[GdATM05] Alexander Gelbukh, Álvaro de Albornoz, and Hugo Terashima-Marín, editors. MICAI 2005: Advances in Artificial Intelligence: 4th Mexican International Conference on Artificial Intelligence, Monterrey, Mexico, November 14– 18, 2005: Proceedings, volume 3789 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29896-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .M48 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3789;
 http://www.springerlink.
com/content/978-3-540-
29896-0.

Guttmann:2011:CAR

[GDG11]  Christian  Guttmann,

Frank Dignum, and Michael Georgeff, editors. Collaborative Agents — Research and Development: International Workshops, CARE@AI09 2009 /CARE@IAT10 2010, Melbourne Australia, December 1, 2009 and Toronto Canada, August 31, 2010, Revised Selected Papers, volume 6066 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22426-1 (print), 3-642-22427-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22427-0.

Gschwind:2012:SCI

[GDGB12] Thomas Gschwind, Flavio De Paoli, Volker Gruhn, and Matthias Book, editors. Software Composition: 11th International Conference, SC 2012, Prague, Czech Republic, May 31 — June 1, 2012. Proceedings, volume 7306 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30563-6 (print), 3-642-30564-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30564-1.

GonzalezHarbour:1999:RST

[GdlP99] Michael González Harbour and Juan A. de la Puente, editors. Reliable software technologies — Ada-Europe ’99:

1999 Ada-Europe International Conference on Reliable Software Technologies, Santander, Spain, June 7–11, 1999: Proceedings, volume 1622 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66093-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1622. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1622.htm;
 http://www.springerlink.
com/content/978-3-540-
66093-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1622.

Ghosh:2007:PRM

[GDP07] Ashish Ghosh, Rajat K.

De, and Sankar K. Pal, editors. Pattern Recognition and Machine Intelligence: Second International Conference, PReMI 2007, Kolkata, India, December 18– 22, 2007. Proceedings, volume 4815 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77045-3 (print), 3-540-77046-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77046-6.

Gangemi:2008:KEP

[GE08] Aldo Gangemi and Jérôme Euzenat, editors. Knowledge Engineering: Practice and Patterns: 16th International Conference, EKAW 2008, Acitrezza, Italy, September 29 — October 2, 2008. Proceedings, volume 5268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87695-2 (print), 3-540-87696-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87696-0.

Gecseg:1981:FCT

[Gec81] F. Gecseg, editor. Fundamentals of computation theory: proceedings of the 1981 International FCT-Conference, Szeged, August 24–28, 1981, volume 117 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10854-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I57 1981. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0117.
htm;   http://www.springerlink.
com/content/978-0-387-
10854-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=117.

Gellersen:1999:HUC

[Gel99] Hans-W. Gellersen, editor. Handheld and ubiquitous computing: First International Symposium, HUC’99, Karlsruhe, Germany, September 27–29, 1999: Proceedings, volume 1707 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66550-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .H35442 1999. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1707.
htm;   http://www.springerlink.
com/content/978-3-540-
66550-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1707.

Gelbukh:2001:CLI

[Gel01] Alexander Gelbukh, editor. Computational linguistics and intelligent text processing: Second International Conference, CICLing 2001, Mexico City, Mexico, February 18–24, 2001: Proceedings, volume 2004 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41687-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.T48 I54 2001; QA76.9.T48 I54 2001 Stacks SEMSTAX-STAX; QA267.A1 L43 no.2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2004.htm;
 http://www.springerlink.
com/content/978-3-540-
41687-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2004.

Gelbukh:2002:CLI

[Gel02] Alexander Gelbukh, editor. Computational linguistics and intelligent text processing: Third International Conference, CICLing 2002, Mexico City, Mexico, February 17–23, 2002: Proceedings, volume 2276 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43219-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.T48 C49 2002; QA76.9.T48 I54 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2276.
htm;   http://www.springerlink.
com/content/978-3-540-
43219-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2276.

Gelbukh:2003:CLI

[Gel03] Alexander Gelbukh, editor. Computational linguistics and intelligent text processing: 4th International Conference, CICLing 2003, Mexico City, Mexico, February 16–22, 2003: Proceedings, volume 2588 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00532-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.T48 I54 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2588.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        00532-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2588.

Also available via the

World Wide Web.

Gelbukh:2004:CLI

[Gel04] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 5th International Conference, CICLing 2004, Seoul, Korea, February 15–21, 2004: Proceedings, volume 2945 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21006-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN P98 .C495 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2945.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21006-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2945;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95558.

Gelbukh:2005:CLI

[Gel05] Alexander Gelbukh, editor. Computational linguistics and intelligent text processing: 6th international conference, CICLING 2005, Mexico City, Mexico, February 13–19, 2005. Proceedings, volume 3406 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24523-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.T48 C49 2005.

URL http://www.springerlink.
com/content/978-3-540-
24523-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3406;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105772.

Gelbukh:2006:CLI

[Gel06] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 7th International Conference, CICLing 2006, Mexico City, Mexico, February 19–25, 2006. Proceedings, volume 3878 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32205-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32205-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3878.

Gelbukh:2007:CLI

[Gel07] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 8th International Conference, CICLing 2007, Mexico City, Mexico, February 18–24, 2007. Proceedings, volume 4394 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70938-X (print), 3-540-70939-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70939-8.

Gelbukh:2008:CLI

[Gel08] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 9th International Conference, CICLing 2008, Haifa, Israel, February 17–23, 2008. Proceedings, volume 4919 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78134-X (print), 3-540-78135-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78135-6.

Gelbukh:2009:CLI

[Gel09] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 10th International Conference, CICLing 2009, Mexico City, Mexico, March 1– 7, 2009. Proceedings, volume 5449 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00381-8 (print), 3-642-00382-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00382-0.

Gelbukh:2010:CLI

[Gel10] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 11th International Conference, CICLing 2010, Iaşi, Romania, March 21–27, 2010. Proceedings, volume 6008 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12115-2 (print), 3-642-12116-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12116-6.

Gelbukh:2011:CLIb

[Gel11a] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 12th International Conference, CICLing 2011, Tokyo, Japan, February 20–26, 2011. Proceedings, Part II, volume 6609 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19436-2 (print), 3-642-19437-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19437-5.

Gelbukh:2011:CLIa

[Gel11b] Alexander F. Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 12th International Conference, CICLing 2011, Tokyo, Japan, February 20–26, 2011. Proceedings, Part I, volume 6608 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19399-4 (print), 3-642-19400-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19400-9.

Gelbukh:2012:CLIa

[Gel12a] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 13th International Conference, CICLing 2012, New Delhi, India, March 11–17, 2012, Proceedings, Part I, volume 7181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28603-8 (print), 3-642-28604-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28604-9.

Gelbukh:2012:CLIb

[Gel12b] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 13th International Conference, CICLing 2012, New Delhi, India, March 11– 17, 2012, Proceedings, Part II, volume 7182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28600-3 (print), 3-642-28601-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28601-8.

Gent:2009:PPC

[Gen09] Ian P. Gent, editor. Principles and Practice of Constraint Programming — CP 2009: 15th International Conference, CP 2009 Lisbon, Portugal, September 20–24, 2009 Proceedings, volume 5732 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04243-0 (print), 3-642-04244-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04244-7.

Gerhard:2004:MAS

[Ger04] Juergen Gerhard. Modular algorithms in symbolic summation and symbolic integration, volume 3218 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-24061-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 224 pp. LCCN QA76.9.A43 G474 2004.

URL http://www.springerlink.
com/content/978-3-540-
24061-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3218.

Geurts:1998:ACA

[Geu98] Frédéric Geurts. Abstract Compositional Analysis of Iterated Relations: A Structural Approach to Complex State Transition Systems, volume 1426 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65506-9. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 280 pp. LCCN QA76.9.A23 G48 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1426.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65506-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1426.

Gunter:1997:TPH

[GF97] Elsa L. Gunter and Amy Felty, editors. Theorem proving in higher order logics: 10th International Conference, TPHOLs ’97, Murray Hill, NJ, USA, August 19–22, 1997: proceedings, volume 1275 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63379-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96I577 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1275.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63379-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1275.

Gedeon:2003:AAA

[GF03] Tamás D. Gedeon and Lance Chun Che Fung, editors. AI 2003: Advances in Artificial Intelligence: 16th Australian Conference on AI, Perth, Australia, December 3–5, 2003: Proceedings, volume 2903 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20646-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A43 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2903.htm;
 http://www.springerlink.
com/content/978-3-540-
20646-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2903;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94701.

Gollmann:2012:ISI

[GF12a] Dieter Gollmann and Felix C. Freiling, editors. Information Security: 15th International Conference, ISC 2012, Passau, Germany, September 19–21, 2012. Proceedings, volume 7483 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33382-6 (print), 3-642-33383-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33383-5.

Groth:2012:PAD

[GF12b] Paul Groth and James Frew, editors. Provenance and annotation of data and processes: 4th International Provenance and Annotation Workshop, IPAW 2012, Santa Barbara, CA, USA, June 19–21, 2012, Revised selected papers, volume 7525 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34221-3, 3-642-34222-1 (e-book). LCCN QA76.9.D3 I58 2012.

                                        URL http://www.springerlink.
                                        com/content/978-3-642-
                                        34221-9;   http://www.
                                        springerlink.com/content/
                                        978-3-642-34222-6.

Gauthier:2000:ITS

[GFV00] Gilles Gauthier, Claude Frasson, and Kurt Van-Lehn, editors. Intelligent tutoring systems: 5th International Conference, ITS 2000, Montréal, Canada, June 19–23, 2000: Proceedings, volume 1839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67655-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1839. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1839.htm;
 http://www.springerlink.
com/content/978-3-540-
67655-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1839.

Geradts:2009:CFT

[GFV09] Zeno J. M. H. Geradts, Katrin Y. Franke, and Cor J. Veenman, editors. Computational Forensics: Third International Workshop, IWCF 2009, The Hague, The Netherlands, August 13–14, 2009. Proceedings, volume 5718 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03520-5 (print), 3-642-03521-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03521-0.

Giles:1998:APS

[GG98] C. Lee Giles and Marco Gori, editors. Adaptive Processing of Sequences and Data Structures: International Summer School on Neural Networks “E. R. Caianiello” Vietri sul Mare, Italy, September 6–13, 1997 tutorial lectures, volume 1387 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64341-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .A36 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1387.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64341-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1387.

Ghelli:2002:DPL

[GG02a] Giorgio Ghelli and Gösta Grahne, editors. Database programming languages: 8th International Workshop, DBPL 2001, Frascati, Italy, September 8– 10, 2001: Revised Papers, volume 2397 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44080-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I585 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2397.htm;
 http://www.springerlink.
com/content/978-3-540-
44080-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2397.

Guigo:2002:ABS

[GG02b] Roderic Guigó and Dan Gusfield, editors. Algorithms in Bioinformatics: Second International Workshop, WABI 2002, Rome, Italy, September 17–21, 2002. Proceedings, volume 2452 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44211-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R858.A2 A44 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2452.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        44211-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2452.

Gabbrielli:2005:LPI

[GG05a] Maurizio Gabbrielli and Gopal Gupta, editors. Logic programming: 21st international conference, ICLP 2005, Barcelona, Spain, October 2–5, 2005, proceedings, volume 3668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29208-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29208-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3668.

Ganter:2005:FCAa

[GG05b] Bernhard Ganter and Robert Godin, editors. Formal concept analysis: third international conference, ICFCA 2005, Lens, France, February 14–18, 2005. Proceedings, volume 3403 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24525-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA171.5 .I24 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24525-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3403;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105806.

Gervasi:2007:CSAa

[GG07a] Osvaldo Gervasi and Marina L. Gavrilova, editors. Computational Science and Its Applications — ICCSA 2007: International Conference, Kuala Lumpur, Malaysia, August 26–29, 2007. Proceedings, Part I, volume 4705 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74468-1 (print), 3-540-74472-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74472-6.

Gervasi:2007:CSAb

[GG07b] Osvaldo Gervasi and Marina L. Gavrilova, editors. Computational Science and Its Applications — ICCSA 2007: International Conference, Kuala Lumpur, Malaysia, August 26–29, 2007. Proceedings, Part II, volume 4706 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74475-4 (print), 3-540-74477-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74477-1.

Gervasi:2007:CSAc

[GG07c] Osvaldo Gervasi and Marina L. Gavrilova, editors. Computational Science and Its Applications — ICCSA 2007: International Conference, Kuala Lumpur, Malaysia, August 26–29, 2007. Proceedings. Part III, volume 4707 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74482-7 (print), 3-540-74484-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74484-9.

Gardner:2009:DPL

[GG09] Philippa Gardner and Floris Geerts, editors. Database Programming Languages: 12th International Symposium, DBPL 2009, Lyon, France, August 23–24, 2009. Proceedings, volume 5708 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03792-5 (print), 3-642-03793-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03793-1.

Gong:2010:PCI

[GG10] Guang Gong and Kishan Chand Gupta, editors. Progress in Cryptology — INDOCRYPT 2010: 11th International Conference on Cryptology in India, Hyderabad, India, December 12–15, 2010. Proceedings, volume 6498 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17400-0 (print), 3-642-17401-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17401-8.

Gaedke:2009:WEI

[GGD09] Martin Gaedke, Michael Grossniklaus, and Oscar Díaz, editors. Web Engineering: 9th International Conference, ICWE 2009 San Sebastián, Spain, June 24– 26, 2009 Proceedings, volume 5648 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02817-9 (print), 3-642-02818-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02818-2.

Gerstner:1997:ANN

[GGHN97] Wulfram Gerstner, Alain Germond, Martin Hasler, and Jean-Daniel Nicoud, editors. Artificial neural networks: ICANN ’97: 7th International Conference, Lausanne, Switzerland, October 8– 10, 1997: proceeedings [sic], volume 1327 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63631-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I56 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1327.htm;
 http://www.springerlink.
com/content/978-3-540-
63631-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1327.

Golomb:2007:SSC

[GGHS07] Solomon W. Golomb, Guang Gong, Tor Helleseth, and Hong-Yeop Song, editors. Sequences, Subsequences, and Consequences: International Workshop, SSC 2007, Los Angeles, CA, USA, May 31 — June 2, 2007, Revised Invited Papers, volume 4893 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77403-3 (print), 3-540-77404-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77404-4.

Gervasi:2005:CSAa

[GGK^{+}05a] Osvaldo Gervasi, Marina L. Gavrilova, Vipin Kumar, Antonio Laganà, Heow Pueh Lee, Youngsong Mun, David Taniar, and Chih Jeng Kenneth Tan, editors. Computational Science and Its Applications: ICCSA 2005: International Conference Singapore, May 9–12, 2005 Proceedings, Part I, volume 3480 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25860-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25860-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3480;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136266.

Gervasi:2005:CSAb

[GGK^{+}05b] Osvaldo Gervasi, Marina L. Gavrilova, Vipin Kumar, Antonio Laganà, Heow Pueh Lee, Youngsong Mun, David Taniar, and Chih Jeng Kenneth Tan, editors. Computational Science and Its Applications: ICCSA 2005: International Conference Singapore, May 9–12, 2005 Proceedings, Part II, volume 3481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25861-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25861-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3481;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136269.

Gervasi:2005:CSAc

[GGK^{+}05c] Osvaldo Gervasi, Marina L. Gavrilova, Vipin Kumar, Antonio Laganà, Heow Pueh Lee, Youngsong Mun, David Taniar, and Chih Jeng Kenneth Tan, editors. Computational Science and Its Applications: ICCSA 2005: International Conference Singapore, May 9–12, 2005 Proceedings, Part III, volume 3482 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25862-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25862-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3482;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136271.

Gervasi:2005:CSAd

[GGK^{+}05d] Osvaldo Gervasi, Marina L. Gavrilova, Vipin Kumar, Antonio Laganà, Heow Pueh Lee, Youngsong Mun, David Taniar, and Chih Jeng Kenneth Tan, editors. Computational Science and Its Applications: ICCSA 2005: International Conference, Singapore, May 9–12, 2005 Proceedings, Part IV, volume 3483 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25863-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25863-6;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3483;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136278.

Gavrilova:2006:PCIa

[GGK06a] Marina Gavrilova, Osvaldo Gervasi, and Vipin Kumar, editors. Pervasive Computing: 4th International Conference, Pervasive 2006, Dublin, Ireland, May 7–10, 2006, Proceedings, volume 3980 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34070-X (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34070-6.

Gavrilova:2006:CSAc

[GGK^{+}06b] Marina Gavrilova, Osvaldo Gervasi, Vipin Kumar, C. J. Kenneth Tan, David Taniar, Antonio Laganá, Youngsong Mun, and Hyunseung Choo, editors. Computational Science and Its Applications — ICCSA 2006: International Conference, Glasgow, UK, May 8–11, 2006, Proceedings, Part III, volume 3982 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34075-0 (print), 3-540-34076-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34076-8.

Gavrilova:2006:CSAb

[GGK^{+}06c] Marina L. Gavrilova, Osvaldo Gervasi, Vipin Kumar, C. J. Kenneth Tan, David Taniar, Antonio Laganá, Youngsong Mun, and Hyunseung Choo, editors. Computational Science and Its Applications — ICCSA 2006: International Conference, Glasgow, UK, May 8–11, 2006. Proceedings, Part II, volume 3981 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34072-6 (print), 3-540-34074-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34074-4.

Gavrilova:2006:CSAd

[GGK^{+}06d] Marina L. Gavrilova, Osvaldo Gervasi, Vipin Kumar, C. J. Kenneth Tan, David Taniar, Antonio Laganá, Youngsong Mun, and Hyunseung Choo, editors. Computational Science and Its Applications — ICCSA 2006: International Conference, Glasgow, UK, May 8–11, 2006, Proceedings, Part IV, volume 3983 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34077-7 (print), 3-540-34078-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34078-2.

Gavrilova:2006:CSAe

[GGK^{+}06e] Marina L. Gavrilova, Osvaldo Gervasi, Vipin Kumar, C. J. Kenneth Tan, David Taniar, Antonio Laganá, Youngsong Mun, and Hyunseung Choo, editors. Computational Science and Its Applications — ICCSA 2006: International Conference, Glasgow, UK, May 8–11, 2006, Proceedings, Part V, volume 3984 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34079-3 (print), 3-540-34080-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34080-5.

Gross:2009:HCIa

[GGK^{+}09a] Tom Gross, Jan Gulliksen, Paula Kotzé, Lars Oestreicher, Philippe Palanque, Raquel Oliveira Prates, and Marco Winckler, editors. Human-Computer Interaction — INTERACT 2009: 12th IFIP TC 13 International Conference, Uppsala, Sweden, August 24– 28, 2009, Proceedings, Part I, volume 5726 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03654-6 (print), 3-642-03655-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03655-2.

Gross:2009:HCIb

[GGK^{+}09b] Tom Gross, Jan Gulliksen, Paula Kotzé, Lars Oestreicher, Philippe Palanque, Raquel Oliveira Prates, and Marco Winckler, editors. Human-Computer Interaction — INTERACT 2009: 12th IFIP TC 13 International Conference, Uppsala, Sweden, August 24– 28, 2009, Proceedings, Part II, volume 5727 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03657-0 (print), 3-642-03658-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03658-3.

Gottlob:1999:CSL

[GGS99] G. Gottlob, E. Grandjean, and Katrin Seyr, editors. Computer science logic: 12th International Workshop, CSL ’98: Annual Conference of the EACSL, Brno, Czech Republic, August 24–28, 1998: Proceedings, volume 1584 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65922-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W624 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1584.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65922-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1584.

Ghose:2009:ACM

[GGS09] Aditya Ghose, Guido Governatori, and Ramakoti Sadananda, editors. Agent Computing and Multi-Agent Systems: 10th Pacific Rim International Conference on Multi-Agents, PRIMA 2007, Bangkok, Thailand, November 21– 23, 2007. Revised Papers, volume 5044 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01638-3 (print), 3-642-01639-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01639-4.

Gleizes:2011:AOS

[GGS11] Marie-Pierre Gleizes and Jorge J. Gomez-Sanz, editors. Agent-Oriented Software Engineering X: 10th International Workshop, AOSE 2009, Budapest, Hungary, May 11–12, 2009, Revised Selected Papers, volume 6038 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19207-6 (print), 3-642-19208-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19208-1.

Gusgen:1992:PCB

[GH92] Hans Werner Gusgen and J. Hertzberg. A perspective of constraint-based reasoning: an introductory tutorial, volume 597 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 0-387-55510-2 (Berlin), 3-540-55510-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 123 pp. LCCN Q340 .G8 1992. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0597.
htm;   http://www.springerlink.
com/content/978-0-387-
55510-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=597.

Grunbacher:1993:FPG

[GH93] H. Grünbacher and Reiner Hartenstein, editors. Field-programmable gate arrays: architecture and tools for rapid prototyping: Second International Workshop on Field-Programmable Logic and Applications, Vienna, Austria, August 31–September 2, 1992: Selected Papers, volume 705 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57091-8 (Berlin), 0-387-57091-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7872.L64 I57 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0705.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        57091-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=705.

Gentzsch:1994:HPCa

[GH94a] Wolfgang Gentzsch and Uwe Harms, editors. High-performance computing and networking: International Conference and Exhibition, Munich, Germany, April 18–20, 1994: proceedings, volume 796 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994). DM96.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0796.htm;
 http://www.springerlink.
com/content/978-3-540-
57980-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=796.

Two volumes.

Gentzsch:1994:HPCb

[GH94b] Wolfgang Gentzsch and Uwe Harms, editors. High-performance computing and networking: International Conference and Exhibition, Munich, Germany, April 18–20, 1994: proceedings, volume 797 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994). DM96.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0797.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        57981-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=797.

Two volumes.

Gorz:1996:KAA

[GH96] Gunther Gorz and Steffen Holldobler, editors. KI-96: advances in artificial intelligence: 20th Annual German Conference on Artificial Intelligence, Dresden, Germany, September 17– 19, 1996: proceedings, volume 1137 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61708-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G47 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1137.htm;
                                         http://www.springerlink.

com/content/978-3-540-
61708-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1137.

Garavel:2003:TAC

[GH03] Hubert Garavel and John Hatcliff, editors. Tools and algorithms for the construction and analysis of systems: 9th International Conference, TACAS 2003, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2003, Warsaw, Poland, April 7–11, 2003: Proceedings, volume 2619 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00898-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T33 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2619.htm;
 http://www.springerlink.
com/content/978-3-540-
00898-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2619.

Also available via the World Wide Web.

Groz:2004:TCS

[GH04] Roland Groz and Robert M. Hierons, editors. Testing of Communicating Systems: 16th IFIP International Conference, Test-Com 2004, Oxford, UK, March 17–19, 2004: Proceedings, volume 2978 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21219-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5101.A1 I3593 2004.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2978.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        21219-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2978;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95741.

Gilbert:2005:FSE

[GH05] Henri Gilbert and Helena Handschuh, editors. Fast Software Encryption: 12th International Workshop, FSE 2005, Paris, France, February 21–23, 2005, Revised Selected Papers, volume 3557 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26541-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F77 2005.

URL http://www.springerlink.
com/content/978-3-540-
26541-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3557.

Gilroy:2006:ISD

[GH06] Stephen W. Gilroy and Michael D. Harrison, editors. Interactive Systems: Design, Specification, and Verification. 12th International Workshop, DSV-IS 2005, Newcastle upon Tyne, UK, July 13–15, 2005. Revised Papers, volume 3941 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34145-5 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34145-1.

Giancarlo:2007:ABI

[GH07a] Raffaele Giancarlo and Sridhar Hannenhalli, editors. Algorithms in Bioinformatics: 7th International Workshop, WABI 2007, Philadelphia, PA, USA, September 8–9, 2007. Proceedings, volume 4645 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74125-9 (print), 3-540-74126-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74126-8.

Grumberg:2007:TAC

[GH07b] Orna Grumberg and Michael Huth, editors. Tools and Algorithms for the Construction and Analysis of Systems: 13th International Conference, TACAS 2007, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2007 Braga, Portugal, March 24 — April 1, 2007. Proceedings, volume 4424 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71208-9 (print), 3-540-71209-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71209-1.

Garrigue:2008:FLP

[GH08] Jacques Garrigue and Manuel V. Hermenegildo, editors. Functional and Logic Programming: 9th International Symposium, FLOPS 2008, Ise, Japan, April 14– 16, 2008. Proceedings, volume 4989 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78968-5 (print), 3-540-78969-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78969-7.

Glinz:2009:REF

[GH09] Martin Glinz and Patrick Heymans, editors. Requirements Engineering: Foundation for Software Quality: 15th International Working Conference, REFSQ 2009 Amsterdam, The Netherlands, June 8–9, 2009 Proceedings, volume 5512 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02049-6 (print), 3-642-02050-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02050-6.

Giesl:2010:ARI

[GH10] Jürgen Giesl and Reiner Hähnle, editors. Automated Reasoning: 5th International Joint Conference, IJCAR 2010, Edinburgh, UK, July 16– 19, 2010. Proceedings, volume 6173 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14202-8 (print), 3-642-14203-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14203-1.

Gill:2012:IAF

[GH12a] Andy Gill and Jurriaan Hage, editors. Implementation and Application of Functional Languages: 23rd International Symposium, IFL 2011, Lawrence, KS, USA, October 3–5, 2011, Revised Selected Papers, volume 7257 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34406-2 (print), 3-642-34407-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34407-7.

Goschka:2012:DAI

[GH12b] Karl Michael Göschka and Seif Haridi, editors. Distributed Applications and Interoperable Systems: 12th IFIP WG 6.1 International Conference, DAIS 2012, Stockholm, Sweden, June 13–16, 2012. Proceedings, volume 7272 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30822-8 (print), 3-642-30823-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30823-9.

Gorton:2006:CBS

[GHC^{+}06] Ian Gorton, George T.

Heineman, Ivica Crnković, Heinz W. Schmidt, Judith A. Stafford, Clemens Szyperski, and Kurt Wallnau, editors. Component-Based Software Engineering: 9th International Symposium, CBSE 2006, Västerås, Sweden, June 29 — July 1, 2006. Proceedings, volume 4063 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35628-2 (print), 3-540-35629-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35629-5.

Greco:2006:RSC

[GHH^{+}06a] Salvatore Greco, Yutaka Hata, Shoji Hirano, Masahiro Inuiguchi, Sadaaki Miyamoto, Hung Son Nguyen, and Roman Słowiński, editors. Rough Sets and Current Trends in Computing: 5th International Conference, RSCTC 2006 Kobe, Japan, November 6–8, 2006 Proceedings, volume 4259 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47693-8 (print), 3-540-49842-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
49842-1.

Grigoriev:2006:CST

[GHH06b] Dima Grigoriev, John Harrison, and Edward A. Hirsch, editors. Computer Science — Theory and Applications: First International Symposium on Computer Science in Russia, Csr 2006, St. Petersburg, Russia, June 8–12. 2006, Proceedings, volume 3967 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34166-8 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34166-6.

Grust:2006:CTD

[GHI^{+}06] Torsten Grust, Hagen Höpfner, Arantza Illarramendi, Stefan Jablonski, Marco Mesiti, Sascha Müller, Paula-Lavinia Patranjan, Kai-Uwe Sattler, Myra Spiliopoulou, and Jef Wijsen, editors. Current Trends in Database Technology — EDBT 2006: EDBT 2006 Workshops PhD, DataX, IIDB, IIHA, ICSNW, QLQP, PIM, PaRMA, and Reactivity on the Web, Munich, Germany, March 26–31, 2006, Revised Selected Papers, volume 4254 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46788-2 (print), 3-540-46790-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46790-8.

Gimelfarb:2012:SSS

[GHI^{+}12] Georgy Gimel’farb, Edwin Hancock, Atsushi Imiya, Arjan Kuijper, Mineichi Kudo, Shinichiro Omachi, Terry Windeatt, and Keiji Yamada, editors. Structural, Syntactic, and Statistical Pattern Recognition: Joint IAPR International Workshop, SSPR&SPR 2012, Hiroshima, Japan, November 7–9, 2012. Proceedings, volume 7626 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34165-9 (print), 3-642-34166-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34166-3.

Gabrys:2006:KBIa

[GHJ06a] Bogdan Gabrys, Robert J.

Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 10th International Conference, KES 2006, Bournemouth, UK, October 9–11, 2006. Proceedings, Part I, volume 4251 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46535-9 (print), 3-540-46536-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46536-2.

Gabrys:2006:KBIb

[GHJ06b] Bogdan Gabrys, Robert J.

Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 10th International Conference, KES 2006, Bournemouth, UK, October 9–11, 2006. Proceedings, Part II, volume 4252 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46537-5 (print), 3-540-46539-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46539-3.

Gabrys:2006:KBIc

[GHJ06c] Bogdan Gabrys, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 10th International Conference, KES 2006, Bournemouth, UK, October 9–11, 2006. Proceedings, Part III, volume 4253 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46542-1 (print), 3-540-46544-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46544-7.

Glaser:1997:PLI

[GHK97] Hugh Glaser, Pieter H. Hartel, and Herbert Kuchen, editors. Programming languages: implementations, logics, and programs: 9th International Symposium, PLILP ’97 including a special track on declarative programming languages in education, Southampton, UK, September 1–3, 1997: proceedings, volume 1292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63398-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7.P56 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1292.
htm;   http://www.springerlink.
com/content/978-3-540-
63398-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1292.

Gurrin:2010:AIR

[GHK^{+}10] Cathal Gurrin, Yulan He, Gabriella Kazai, Udo Kruschwitz, Suzanne Little, Thomas Roelleke, Stefan Rüger, and Keith van Rijsbergen, editors. Advances in Information Retrieval: 32nd European Conference on IR Research, ECIR 2010, Milton Keynes, UK, March 28–31, 2010.Proceedings, volume 5993 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12274-4 (print), 3-642-12275-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12275-0.

Gallagher:2005:IDE

[GHM05] Marcus Gallagher, James Hogan, and Frederic Maire, editors. Intelligent data engineering and automated learning — IDEAL 2005: 6th international conference, Brisbane, Australia, July 6–8, 2005: proceedings, volume 3578 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26972-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55864 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        26972-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3578.

Grelck:2005:IAF

[GHMT05] Clemens Grelck, Frank Huch, Greg J. Michaelson, and Phil Trinder, editors. Implementation and application of functional languages: 16th international workshop, IFL 2004, Lübeck, Germany, September 8–10, 2004. Revised selected papers, volume 3474 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26094-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26094-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3474;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136772.

Gunopulos:2011:MLKa

[GHMV11a] Dimitrios Gunopulos, Thomas Hofmann, Donato Malerba, and Michalis Vazirgiannis, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2011, Athens, Greece, September 5–9, 2011. Proceedings, Part I, volume 6911 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23779-7 (print), 3-642-23780-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23780-5.

Gunopulos:2011:MLKb

[GHMV11b] Dimitrios Gunopulos, Thomas Hofmann, Donato Malerba, and Michalis Vazirgiannis, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2011, Athens, Greece, September 5–9, 2011, Proceedings, Part II, volume 6912 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23782-7 (print), 3-642-23783-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23783-6.

Gunopulos:2011:MLKc

[GHMV11c] Dimitrios Gunopulos, Thomas Hofmann, Donato Malerba, and Michalis Vazirgiannis, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2011, Athens, Greece, September 5–9, 2011, Proceedings, Part III, volume 6913 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23807-6 (print), 3-642-23808-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23808-6.

Ghosh:2010:MSE

[Gho10] Sudipto Ghosh, editor. Models in Software Engineering: Workshops and Symposia at MODELS 2009, Denver, CO, USA, October 4–9, 2009, Reports and Revised Selected Papers, volume 6002 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12260-4 (print), 3-642-12261-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12261-3.

Gulliksen:2008:EIS

[GHP^{+}08] Jan Gulliksen, Morton Borup Harning, Philippe Palanque, Gerrit C. van der Veer, and Janet Wesson, editors. Engineering Interactive Systems: EIS 2007 Joint Working Conferences, EHCI 2007, DSV-IS 2007, HCSE 2007, Salamanca, Spain, March 22– 24, 2007. Selected Papers, volume 4940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92697-6 (print), 3-540-92698-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92698-6.

Governatori:2009:RIA

[GHP09] Guido Governatori, John Hall, and Adrian Paschke, editors. Rule Interchange and Applications: International Symposium, RuleML 2009, Las Vegas, Nevada, USA, November 5–7, 2009. Proceedings, volume 5858 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04984-2 (print), 3-642-04985-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04985-9.

Gheorghe:2011:MCI

[GHP^{+}11] Marian Gheorghe, Thomas Hinze, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa, editors. Membrane Computing: 11th International Conference, CMC 2010, Jena, Germany, August 24–27, 2010. Revised Selected Papers, volume 6501 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18122-8 (print), 3-642-18123-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18123-8.

Gawlick:1989:HPT

[GHR89] D. Gawlick, M. Haynie, and A. Reuter, editors. High performance transaction systems: 2nd International Workshop, Asilomar Conference Center, Pacific Grove, CA, USA, September 28–30, 1987: proceedings, volume 359 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51085-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.545 .I571 1987. US$26.90. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0359.htm;
 http://www.springerlink.
com/content/978-0-387-
51085-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=359.

Proceedings of the Second International Workshop on High Performance Transaction Systems.

Giorgini:2004:AOI

[GHSW04] Paolo Giorgini, Brian Henderson-Sellers, and Michael Winikoff, editors. Agent-Oriented Information Systems: 5th International Biconference Workshop, AOIS 2003, Melbourne, Australia, July 14, 2003 and Chicago, IL, USA, October 13, 2003: Revised Selected Papers, volume 3030 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22127-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 A48 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3030.
htm;   http://www.springerlink.
com/content/978-3-540-
22127-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3030;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98189.

Gong:2006:STA

[GHSY06] Guang Gong, Tor Helleseth, Hong-Yeop Song, and Kyeongcheol Yang, editors. Sequences and Their Applications — SETA 2006: 4th International Conference Beijing, China, September 24–28, 2006 Proceedings, volume 4086 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44523-4 (print), 3-540-44524-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
44524-1.

Guerra:2003:MMP

[GI03] Concettina Guerra and Sorin Istrail, editors. Mathematical Methods for Protein Structure Analysis and Design: C.I.M.E Summer School, Martinal Franca, Italy, July 9–15, 2000: Advanced Lectures, volume 2666 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40104-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP551 .M335 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2666.htm;
 http://www.springerlink.
com/content/978-3-540-
40104-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2666.

Gianni:1989:SAC

[Gia89] P. Gianni, editor. Symbolic and algebraic computation: International Symposium ISSAC ’88, Rome, Italy, July 4– 8, 1988: proceedings, volume 358 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51084-2 (USA).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .I571 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0358.htm;
 http://www.springerlink.
com/content/978-0-387-
51084-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=358.

Conference held jointly with AAECC-6.

Giacobazzi:2004:SAI

[Gia04] Roberto Giacobazzi, editor. Static analysis: 11th international symposium, SAS 2004, Verona, Italy, August 26–28, 2004. proceedings, volume 3148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22791-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
22791-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3148;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99688.

Giacobini:2007:AEC

[Gia07] Mario Giacobini, editor. Applications of Evolutionary Computing: EvoWorkshops 2007: Evo-CoMnet, EvoFIN, EvoIASP,EvoINTERACTION, EvoMUSART, EvoS-TOC and EvoTransLog. Proceedings, volume 4448 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71804-4 (print), 3-540-71805-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71805-5.

Gibbs:1989:SEE

[Gib89] Norman E. Gibbs, editor. Software engineering education: SEI conference 1989, Pittsburgh, Pennsylvania, USA, July 18–21, 1989 proceedings, volume 376 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-97090-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S6531 1989. DM47.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/

0558/tocs/t0376.htm;
http://www.springerlink.
com/content/978-0-387-
97090-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=376.

Consists of papers from the Third SEI Conference on Software Engineering Education and the SEI Workshop on an Undergraduate Software Engineering Curriculum.

Gibbons:2012:GIP

[Gib12] Jeremy Gibbons, editor. Generic and Indexed Programming: International Spring School, SSGIP 2010, Oxford, UK, March 22–26, 2010, Revised Lectures, volume 7470 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32201-8 (print), 3-642-32202-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32202-0.

Giesl:2005:TRA

[Gie05] Jürgen Giesl, editor.

Term Rewriting and Applications: 16th International Conference, RTA 2005, Nara, Japan, April 19–21, 2005: Proceedings, volume 3467 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25596-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25596-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3467;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b135673.

Giese:2008:MSE

[Gie08] Holger Giese, editor. Models in Software Engineering: Workshops and Symposia at MoDELS 2007, Nashville, TN, USA, September 30 — October 5, 2007, Reports and Revised Selected Papers, volume 5002 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69069-7 (print), 3-540-69073-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
69073-3.

Giese:2010:ACS

[Gie10] Holger Giese, editor. Architecting Critical Systems: First International Symposium, ISARCS 2010, Prague, Czech Republic, June 23–25, 2010 Proceedings, volume 6150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13555-2 (print), 3-642-13556-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13556-9.

Gilbert:2010:ACE

[Gil10] Henri Gilbert, editor.

Advances in cryptology — EUROCRYPT 2010: 29th Annual International Conference on the Theory and Applications of Cryptographic Techniques, French Riviera, May 30–June 3, 2010. Proceedings, volume 6110 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-13190-5, 3-642-13189-1. LCCN QA76.9.A25 E97 2010. URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13189-9;   http://www.
                                        springerlink.com/content/
                                        978-3-642-13190-5.

Girard:1999:TLC

[Gir99] Jean-Yves Girard, editor. Typed lambda calculi and applications: 4th International Conference, TLCA’99, L’Aquila, Italy, April 7–9, 1999: Proceedings, volume 1581 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65763-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.5.I58 1999. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1581.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65763-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1581.

Giunchiglia:1998:AII

[Giu98] Fausto Giunchiglia, editor. Artificial intelligence: 8th International Conference, AIMSA’98, Sozopol, Bulgaria, September 21–23, 1998: Proceedings, volume 1480 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64993-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1480. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1480.htm;
 http://www.springerlink.
com/content/978-3-540-
64993-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1480.

Ganzinger:1986:PDO

[GJ86] H. Ganzinger and Neil D.

Jones, editors. Programs as data objects: proceedings of a workshop, Copenhagen, Denmark, October 17–19, 1985, volume 217 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16446-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .P761 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0217.htm;
 http://www.springerlink.
com/content/978-0-387-
16446-5;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=217.

Gaudel:1993:TTP

[GJ93] M.-C. Gaudel and Jean-Pierre Jouannaud, editors. TAPSOFT ’93: theory and practice of software development: 4th International Joint Conference CAAP/FASE, Orsay, France, April 13–17, 1993: proceedings, volume 668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56610-4 (Berlin), 0-387-56610-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I58 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0668.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        56610-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=668.

Gao:2009:AAI

[GJ09] Yong Gao and Nathalie Japkowicz, editors. Advances in Artificial Intelligence: 22nd Canadian Conference on Artificial Intelligence, Canadian AI 2009 Kelowna, Canada, May 25–27, 2009 Proceedings, volume 5549 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01817-3 (print), 3-642-01818-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01818-3.

Goel:2010:DRI

[GJN10] Ashok K. Goel, Mateja Jamnik, and N. Hari Narayanan, editors. Diagrammatic Representation and Inference: 6th International Conference, Diagrams 2010, Portland, OR, USA, August 9–11, 2010. Proceedings, volume 6170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14599-X (print), 3-642-14600-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14600-8.

Goel:2008:ARC

[GJRR08] Ashish Goel, Klaus Jansen, José D. P. Rolim, and Ronitt Rubinfeld, editors. Approximation, Randomization and Combinatorial Optimization. Algorithms and Techniques: 11th International Workshop, AP-PROX 2008, and 12th International Workshop, RANDOM 2008, Boston, MA, USA, August 25– 27, 2008. Proceedings, volume 5171 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85362-6 (print), 3-540-85363-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85363-3.

Goldberg:2011:ARC

[GJRR11] Leslie Ann Goldberg, Klaus Jansen, R. Ravi, and José D. P. Rolim, editors. Approximation, Randomization, and Combinatorial Optimization. Algorithms and Techniques: 14th International Workshop, AP-PROX 2011, and 15th International Workshop, RANDOM 2011, Princeton, NJ, USA, August 17–19, 2011. Proceedings, volume 6845 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22934-4 (print), 3-642-22935-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22935-0.

Gupta:2012:ARC

[GJRS12] Anupam Gupta, Klaus Jansen, José Rolim, and Rocco Servedio, editors. Approximation, Randomization, and Combinatorial Optimization. Algorithms and Techniques: 15th International Workshop, AP-PROX 2012, and 16th International Workshop, RANDOM 2012, Cambridge, MA, USA, August 15–17, 2012. Proceedings, volume 7408 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32511-4 (print), 3-642-32512-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32512-0.

Goemans:2001:ARC

[GJRT01] Michel Goemans, K. Jansen, J. D. P. Rolim, and L. Trevisan, editors. Approximation, randomization, and combinatorial optimization: algorithms and techniques: 4th International Workshop on Approximation Algorithms for Combinatorial Optimization Problems, APPROX 2001 and 5th International Workshop on Randomization and Approximation Techniques in Computer Science, RAN-DOM 2001, Berkeley, CA, USA, August 2001: Proceedings, volume 2129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42470-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I643 2001; QA267.A1 L43 no.2129. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2129.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42470-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2129.

Gray:2008:SDI

[GJS08] Alex Gray, Keith Jeffery, and Jianhua Shao, editors. Sharing Data, Information and Knowledge: 25th British National Conference on Databases, BNCOD 25, Cardiff, UK, July 7– 10, 2008. Proceedings, volume 5071 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70503-1 (print), 3-540-70504-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70504-8.

Gavalda:2003:ALT

[GJT03] Ricard Gavaldà, Klaus P.

Jantke, and Eiji Takimoto, editors. Algorithmic Learning Theory: 14th International Conference, ALT 2003, Sapporo, Japan, October 2003: Proceedings, volume 2842 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20291-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2842.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20291-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2842;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b14273.

Gunsel:2006:MCR

[GJTS06] Bilge Gunsel, Anil K. Jain, A. Murat Tekalp, and Bülent Sankur, editors. Multimedia Content Representation, Classification and Security: International Workshop, MRCS 2006, Istanbul, Turkey, September 11– 13, 2006. Proceedings, volume 4105 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39392-7 (print), 3-540-39393-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        39393-1.

Gilbert:1990:SSW

[GK90] J. R. Gilbert and R. Karlsson, editors. SWAT 90: 2nd Scandinavian Workshop on Algorithm Theory, Berger, Sweden [i.e., Norway], July 11–14, 1990: proceedings, volume 447 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 3-540-52846-6, 0-387-52846-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .S28 1990. DM57.50.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0447.
htm;   http://www.springerlink.
com/content/978-3-540-
52846-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=447.

“Presented at the Second Scandinavian Workshop on Algorithm Theory held July 11–14, 1990 in Bergen, Norway”– Foreword.

Goble:1995:ADB

[GK95] Carole Goble and John Keane, editors. Advances in databases: 13th British National Conference on Databases, BNCOD 13 Manchester, United Kingdom, July 12–14, 1995: proceedings, volume 940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60100-7 (Berlin: softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.940.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0940.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60100-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=940.

Gogolla:2001:UUM

[GK01] Martin Gogolla and Cris Kobryn, editors. UML 2001: the unified modeling language: modeling languages, concepts, and tools: 4th International Conference, Toronto, Canada, October 1–5, 2001: Proceedings, volume 2185 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42667-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2185. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2185.htm;

http://www.springerlink.
com/content/978-3-540-
42667-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2185.

Goodrich:2002:GDI

[GK02] Michael T. Goodrich and Stephen G. Kobourov, editors. Graph Drawing: 10th International Symposium, GD 2002, Irvine, CA, USA, August 26–28, 2002. Revised Papers, volume 2528 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00158-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2528.
htm;   http://www.springerlink.
com/content/978-3-540-
00158-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2528.

Also available via the World Wide Web.

Gerndt:2006:HPC

[GK06] Michael Gerndt and Dieter Kranzlmüller, editors. High Performance Computing and Communications: Second International Conference, HPCC 2006, Munich, Germany, September 13– 15, 2006. Proceedings, volume 4208 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39368-4 (print), 3-540-39372-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
39372-6.

Gradel:2009:CSL

[GK09] Erich Grädel and Reinhard Kahle, editors. Computer Science Logic: 23rd international Workshop, CSL 2009, 18th Annual Conference of the EACSL, Coimbra, Portugal, September 7– 11, 2009. Proceedings, volume 5771 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04026-8 (print), 3-642-04027-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04027-6.

Glimm:2012:KAA

[GK12] Birte Glimm and Antonio Krüger, editors. KI 2012: Advances in Artificial Intelligence: 35th Annual German Conference on AI, Saarbrücken, Germany, September 24– 27, 2012. Proceedings, volume 7526 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33346-X (print), 3-642-33347-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33347-7.

Geffert:2008:STP

[GKB^{+}08] Viliam Geffert, Juhani Karhumäki, Alberto Bertoni, Bart Preneel, Pavol Návrat, and Mária Bieliková, editors. SOFSEM 2008: Theory and Practice of Computer Science: 34th Conference on Current Trends in Theory and Practice of Computer Science, Nový Smokovec, Slovakia, January 19– 25, 2008. Proceedings, volume 4910 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77565-X (print), 3-540-77566-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77566-9.

Ge:2012:SRI

[GKC^{+}12] Shuzhi Sam Ge, Oussama Khatib, John-John Cabibihan, Reid Simmons, and Mary-Anne Williams, editors. Social Robotics: 4th International Conference, ICSR 2012, Chengdu, China, October 29–31, 2012. Proceedings, volume 7621 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34102-0 (print), 3-642-34103-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34103-8.

Grumberg:2009:LFN

[GKKW09] Orna Grumberg, Michael Kaminski, Shmuel Katz, and Shuly Wintner, editors. Languages: From Formal to Natural: Essays Dedicated to Nissim Francez on the Occasion of His 65th Birthday, volume 5533 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01747-9 (print), 3-642-01748-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01748-3.

Gal:2009:PPW

[GKL09] Cecilia S. Gal, Paul B.

Kantor, and Michael E. Lesk, editors. Protecting Persons While Protecting the People: Second Annual Workshop on Information Privacy and National Security, ISIPS 2008, New Brunswick, NJ, USA, May 12, 2008. Revised Selected Papers, volume 5661 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10232-8 (print), 3-642-10233-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10233-2.

Giese:2011:MBE

[GKL^{+}11] Holger Giese, Gabor Karsai, Edward Lee, Bernhard Rumpe, and Bernhard Schätz, editors. Model-Based Engineering of Embedded Real-Time Systems: International Dagstuhl Workshop, Dagstuhl Castle, Germany, November 4–9, 2007. Revised Selected Papers, volume 6100 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16276-2 (print), 3-642-16277-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16277-0.

Grundspenkis:2010:ADI

[GKMN10] Janis Grundspenkis, Marite Kirikova, Yannis Manolopoulos, and Leonids Novickis, editors. Advances in Databases and Information Systems: Associated Workshops and Doctoral Consortium of the 13th East European Conference, ADBIS 2009, Riga, Latvia, September 7–10, 2009. Revised Selected Papers, volume 5968 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12081-4 (print), 3-642-12082-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12082-4.

Gerdt:2010:CAS

[GKMV10] Vladimir P. Gerdt, Wolfram Koepf, Ernst W. Mayr, and Evgenii V. Vorozhtsov, editors. Computer Algebra in Scientific Computing: 12th International Workshop, CASC 2010, Tsakhkadzor, Armenia, September 6–12, 2010. Proceedings, volume 6244 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15273-2 (print), 3-642-15274-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15274-0.

Gerdt:2011:CAS

[GKMV11] Vladimir P. Gerdt, Wolfram Koepf, Ernst W. Mayr, and Evgenii V. Vorozhtsov, editors. Computer Algebra in Scientific Computing: 13th International Workshop, CASC 2011, Kassel, Germany, September 5– 9, 2011. Proceedings, volume 6885 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23567-0 (print), 3-642-23568-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23568-9.

Gerdt:2012:CAS

[GKMV12] Vladimir P. Gerdt, Wolfram Koepf, Ernst W. Mayr, and Evgenii V. Vorozhtsov, editors. Computer Algebra in Scientific Computing: 14th International Workshop, CASC 2012, Maribor, Slovenia, September 3– 6, 2012. Proceedings, volume 7442 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32972-1 (print), 3-642-32973-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32973-9.

Gunter:2003:KAA

[GKN03] Andreas Günter, Rudolf Kruse, and Bernd Neumann, editors. KI 2003: Advances in Artificial Intelligence: 26th Annual German Conference on AI, KI 2003, Hamburg, Germany, September 15–18, 2003: Proceedings, volume 2821 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20059-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G466 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2821.htm;
http://www.springerlink.
com/content/978-3-540-
20059-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2821;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13477.

Gurevich:2000:ASM

[GKOT00] Yuri Gurevich, Philipp W.

Kutter, Martin Odersky, and Lothar Thiele, editors. Abstract State Machines: Theory and Applications: International Workshop, ASM 2000, Monte Verità , Switzerland, March 19–24, 2000. Proceedings, volume 1912 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67959-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A25 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1912.htm;
 http://www.springerlink.
com/content/978-3-540-
67959-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1912.

Gorodetsky:2005:CNS

[GKS05] Vladimir Gorodetsky, Igor Kotenko, and Victor Skormin, editors. Computer network security: Third International Workshop on Mathematical Methods, Models, and Architectures for Computer Network Security, MMM-ACNS 2005, St. Petersburg, Russia, September 25–27, 2005: proceedings, volume 3685 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29113-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I565 2005.

URL http://www.springerlink.
com/content/978-3-540-
29113-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3685.

Geraets:2007:AMR

[GKS^{+}07] Frank Geraets, Leo Kroon, Anita Schoebel, Dorothea Wagner, and Christos D. Zaroliagis, editors. Algorithmic Methods for Railway Optimization: International Dagstuhl Workshop, Dagstuhl Castle, Germany, June 20– 25, 2004, 4th International Workshop, AT-MOS 2004, Bergen, Norway, September 16–17, 2004, Revised Selected Papers, volume 4359 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74245-X (print), 3-540-74247-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74247-0.

Geva:2012:FRC

[GKS12] Shlomo Geva, Jaap Kamps, and Ralf Schenkel, editors. Focused Retrieval of Content and Structure: 10th International Workshop of the Initiative for the Evaluation of XML Retrieval, INEX 2011, Saarbrücken, Germany, December 12–14, 2011, Revised Selected Papers, volume 7424 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35733-4 (print), 3-642-35734-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35734-3.

Geva:2011:CEF

[GKST11] Shlomo Geva, Jaap Kamps, Ralf Schenkel, and Andrew Trotman, editors. Comparative Evaluation of Focused Retrieval: 9th International Workshop of the Inititative for the Evaluation of XML Retrieval, INEX 2010, Vugh, The Netherlands, December 13–15, 2010, Revised Selected Papers, volume 6932 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23576-X (print), 3-642-23577-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23577-1.

Geva:2009:AFR

[GKT09] Shlomo Geva, Jaap Kamps, and Andrew Trotman, editors. Advances in Focused Retrieval: 7th International Workshop of the Initiative for the Evaluation of XML Retrieval, INEX 2008, Dagstuhl Castle, Germany, December 15–18, 2008. Revised and Selected Papers, volume 5631 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03760-7 (print), 3-642-03761-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03761-0.

Geva:2010:FRE

[GKT10] Shlomo Geva, Jaap Kamps, and Andrew Trotman, editors. Focused Retrieval and Evaluation: 8th International Workshop of the Initiative for the Evaluation of XML Retrieval, INEX 2009, Brisbane, Australia, December 7– 9, 2009, Revised and Selected Papers, volume 6203 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14555-8 (print), 3-642-14556-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14556-8.

Gollmann:2002:CSE

[GKW02] Dieter Gollmann, Günter Karjoth, and Michael Waidner, editors. Computer security — ES-ORICS 2002: 7th European Symposium on Research in Computer Security, Zürich, Switzerland, October 14–16, 2002: Proceedings, volume 2502 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44345-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E98 2002. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t2502.htm;
 http://www.springerlink.
com/content/978-3-540-
44345-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2502.

Glowinski:1974:CMAa

[GL74a] Roland Glowinski and Jacques Louis Lions, editors. Computing Methods in Applied Sciences and Engineering. International Symposium, Versailles, December 17–21, 1973. Part I, volume 10 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CO-DEN LNCSD9. ISBN 0-387-06768-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA300 .C631 v. 1– 2 (1974); QA300.I49 1973. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0010.htm;
 http://www.springerlink.
com/content/978-0-387-
06768-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=10.

English or French. Organized by IRIA LA-BORIA, Institut de recherche d’informatique et d’automatique.

Glowinski:1974:CMAb

[GL74b] Roland Glowinski and Jacques Louis Lions, editors. Computing Methods in Applied Sciences and Engineering. International Symposium, Versailles, December 17–21, 1973. Part II, volume 11 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 3-540-06769-8, 0-387-06769-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA300 .C631 v. 1–2 (1974). URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0011.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        06769-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=11.

English or French. Organized by IRIA LA-BORIA, Institut de recherche d’informatique et d’automatique.

Gray:1992:ADS

[GL92] Peter M. D. Gray and Robert Lucas, editors. Advanced database systems: 10th British National Conference on Databases, BNCOD 10, Aberdeen, Scotland, July 6–8, 1992: proceedings, volume 618 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55693-1 (Berlin), 0-387-55693-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B75 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0618.htm;
 http://www.springerlink.
com/content/978-3-540-
55693-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=618.

Garlan:1997:CLM

[GL97] David Garlan and Daniel Le Métayer, editors. Coordination languages and models: second international conference, COORDINATION ’97, Berlin, Germany, September 1–3, 1997: proceedings, volume 1282 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63383-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.C695 1997. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1282.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63383-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1282.

Griffin:2003:SSM

[GL03] Lewis D. Griffin and Martin Lillholm, editors. Scale space methods in computer vision: 4th International Conference, Scale Space 2003, Isle of Skye, UK, June 10–12, 2003: Proceedings, volume 2695 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40368-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .S38 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2695.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40368-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2695.

Gluck:2005:GPC

[GL05] Robert Glück and Michael Lowry, editors. Generative programming and component engineering: 4th international conference, GPCE 2005, Tallinn, Estonia, September 29–October 1, 2005, proceedings, volume 3676 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29138-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29138-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3676.

Greco:2008:SUM

[GL08] Sergio Greco and Thomas Lukasiewicz, editors. Scalable Uncertainty Management: Second International Conference, SUM 2008, Naples, Italy, October 1–3, 2008. Proceedings, volume 5291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87992-7 (print), 3-540-87993-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87993-0.

Gastin:2010:CCT

[GL10a] Paul Gastin and François Laroussinie, editors. CON-CUR 2010 — Concurrency Theory: 21th International Conference, CONCUR 2010, Paris, France, August 31–September 3, 2010. Proceedings, volume 6269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15374-7 (print), 3-642-15375-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15375-4.

Gertz:2010:SSD

[GL10b] Michael Gertz and Bertram Ludäscher, editors. Scientific and Statistical Database Management: 22nd International Conference, SSDBM 2010, Heidelberg, Germany, June 30–July 2, 2010. Proceedings, volume 6187 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13817-9 (print), 3-642-13818-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13818-8.

Gong:2011:WISa

[GLC^{+}11a] Zhiguo Gong, Xiangfeng Luo, Junjie Chen, Jingsheng Lei, and Fu Lee Wang, editors. Web Information Systems and Mining: International Conference, WISM 2011, Taiyuan, China, September 24–25, 2011, Proceedings, Part I, volume 6987 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23970-6 (print), 3-642-23971-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23971-7.

Gong:2011:WISb

[GLC^{+}11b] Zhiguo Gong, Xiangfeng Luo, Junjie Chen, Jingsheng Lei, and Fu Lee Wang, editors. Web Information Systems and Mining: International Conference, WISM 2011, Taiyuan, China, September 24–25, 2011, Proceedings, Part II, volume 6988 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23981-1 (print), 3-642-23982-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23982-3.

Ge:2010:SRS

[GLCT10] Shuzhi Sam Ge, Haizhou Li, John-John Cabibihan, and Yeow Kee Tan, editors. Social Robotics: Second International Conference on Social Robotics, ICSR 2010, Singapore, November 23–24, 2010. Proceedings, volume 6414 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17247-4 (print), 3-642-17248-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17248-9.

Gollmann:2010:SCR

[GLIC10] Dieter Gollmann, Jean-Louis Lanet, and Julien Iguchi-Cartigny, editors. Smart card research and advanced application: 9th IFIP WG 8.8/11.2 International Conference, CARDIS 2010, Passau, Germany, April 14–16, 2010: proceedings, volume 6035 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-12509-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.S62 C36 2010. URL http:

//www.springerlink.
com/content/978-3-642-
12509-6.

Gottlob:1993:CLP

[GLM93] G. Gottlob, Alexander Leitsch, and Daniele Mundici, editors. Computational logic and proof theory: Third Kurt Godel Colloquium, KGC’93, Brno, Czech Republic, August 1993: proceedings, volume 713 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57184-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 K87 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0713.
htm;   http://www.springerlink.
com/content/978-0-387-
57184-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=713.

Sponsored by Kurt Godel

Society.

Gottlob:1997:CLP

[GLM97] G. (Georg) Gottlob, Alexander Leitsch, and Daniele Mundici, editors. Computational logic and proof theory: 5th Kurt Godel Colloquium, KGC ’97, Vienna, Austria, August 25–29, 1997: proceedings, volume 1289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63385-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96K87 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1289.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63385-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1289.

Girgin:2008:RAR

[GLM^{+}08] Sertan Girgin, Manuel Loth, Rémi Munos, Philippe Preux, and Daniil Ryabko, editors. Recent Advances in Reinforcement Learning: 8th European Workshop, EWRL 2008, Villeneuve d’Ascq, France, June 30–July 3, 2008, Revised and Selected Papers, volume 5323 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89721-6 (print), 3-540-89722-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89722-4.

Garay:2007:ISI

[GLMP07] Juan A. Garay, Arjen K. Lenstra, Masahiro Mambo, and René Peralta, editors. Information Security: 10th International Conference, ISC 2007, Valparaíso, Chile, October 9–12, 2007. Proceedings, volume 4779 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75495-4 (print), 3-540-75496-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75496-1.

Gore:2001:ARF

[GLN01] Rajeev Goré, Alexander Leitsch, and Tobias Nipkow, editors. Automated reasoning: First International Joint Conference, IJCAR 2001, Siena, Italy, June 18–23, 2001: Proceedings, volume 2083 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42254-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I38 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2083.htm;
 http://www.springerlink.
com/content/978-3-540-
42254-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2083.

Gelfond:1999:LPN

[GLP99] Michael Gelfond, Nicola Leone, and Gerald Pfeifer, editors. Logic programming and nonmonotonic reasoning: 5th International Conference, LPNMR ’99, El Paso, Texas, USA, December 2–4, 1999: Proceedings, volume 1730 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66749-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1730. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1730.htm;
http://www.springerlink.
com/content/978-3-540-
66749-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1730.

Ganascia:2012:DSI

[GLP12] Jean-Gabriel Ganascia, Philippe Lenca, and Jean-Marc Petit, editors. Discovery Science: 15th International Conference, DS 2012, Lyon, France, October 29–31, 2012. Proceedings, volume 7569 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33491-1 (print), 3-642-33492-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33492-4.

Lee:2012:DSAa

[gLPZ^{+}12a] Sang goo Lee, Zhiyong Peng, Xiaofang Zhou, Yang-Sae Moon, Rainer Unland, and Jaesoo Yoo, editors. Database Systems for Advanced Applications: 17th International Conference, DASFAA 2012, Busan, South Korea, April 15– 19, 2012, Proceedings, Part I, volume 7238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29037-X (print), 3-642-29038-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29038-1.

Lee:2012:DSAb

[gLPZ^{+}12b] Sang goo Lee, Zhiyong Peng, Xiaofang Zhou, Yang-Sae Moon, Rainer Unland, and Jaesoo Yoo, editors. Database Systems for Advanced Applications: 17th International Conference, DASFAA 2012, Busan, South Korea, April 15– 19, 2012, Proceedings, Part II, volume 7239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29034-5 (print), 3-642-29035-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29035-0.

Gorodetsky:2005:AIS

[GLS05] Vladimir Gorodetsky, Jiming Liu, and Victor A. Skormin, editors. Autonomous intelligent systems: agents and data mining: International Workshop, AIS-ADM 2005 St. Petersburg, Russia, June 6–8, 2005 Proceedings, volume 3505 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26164-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26164-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3505;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136855.

Gao:2010:DLT

[GLSY10] Yuan Gao, Hanlin Lu, Shinnosuke Seki, and Sheng Yu, editors. Developments in Language Theory: 14th International Conference, DLT 2010, London, ON, Canada, August 17– 20, 2010. Proceedings, volume 6224 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14454-3 (print), 3-642-14455-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14455-4.

Gasevic:2009:SLE

[GLV09] Dragan Gašević, Ralf Lämmel, and Eric Van Wyk, editors. Software Language Engineering: First International Conference, SLE 2008, Toulouse, France, September 29–30, 2008. Revised Selected Papers, volume 5452 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00433-4 (print), 3-642-00434-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00434-6.

Grabowski:1988:ALP

[GLW88] J. Grabowski, Pierre Lescanne, and Wolfgang Wechler, editors. Algebraic and logic programming: international workshop, Gaussig, GDR, November 14–18, 1988: proceedings, volume 343 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-50667-5 (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .A45831 1989. DM40.50. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0343.htm;
http://www.springerlink.
com/content/978-0-387-
50667-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=343.

George:2007:DMD

[GLW07] Chris W. George, Zhiming Liu, and Jim Woodcock, editors. Domain Modeling and the Duration Calculus: International Training School, Shanghai, China, September 17–21. 2007, Advanced Lectures, volume 4710 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74963-2 (print), 3-540-74964-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74964-6.

Gao:2012:WAI

[GLW^{+}12] Hong Gao, Lipyeow Lim, Wei Wang, Chuan Li, and Lei Chen, editors. Web-Age Information Management: 13th International Conference, WAIM 2012, Harbin, China, August 18–20, 2012. Proceedings, volume 7418 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32280-8 (print), 3-642-32281-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32281-5.

Garcia:2003:SEL

[GLZ^{+}03] Alessandro Garcia, Carlos Lucena, Franco Zambonelli, Andrea Omicini, and Jaelson Castro, editors. Software Engineering for Large-Scale Multi-Agent Systems: Research Issues and Practical Applications, volume 2603 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-08772-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 284 pp. LCCN QA76.758 .S64635 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2603.
htm;   http://www.springerlink.
com/content/978-3-540-
08772-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2603.

Also available via the World Wide Web.

Gavalda:2009:ALT

[GLZZ09] Ricard Gavaldà, Gábor Lugosi, Thomas Zeugmann, and Sandra Zilles, editors. Algorithmic Learning Theory: 20th International Conference, ALT 2009, Porto, Portugal, October 3– 5, 2009. Proceedings, volume 5809 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04413-1 (print), 3-642-04414-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04414-4.

Ghezzi:1989:EES

[GM89] Carlo Ghezzi and John McDermid, editors. ESEC ’89: 2nd European Software Engineering Conference, University of Warwick, Coventry, UK, September, 11–15, 1989: proceedings, volume 387 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51635-2 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .E971 1989. DM63.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0387.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51635-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=387.

Goodrich:1999:AEE

[GM99] Michael T. Goodrich and Catherine C. McGeoch, editors. Algorithm engineering and experimentation: International Workshop ALENEX ’99, Baltimore, MD, USA, January 15–16, 1999: Selected Papers, volume 1619 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66227-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A42 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1619.htm;
 http://www.springerlink.
com/content/978-3-540-
66227-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1619.

Ganter:2000:CSL

[GM00] Bernhard Ganter and Guy W. Mineau, editors. Conceptual structures: logical, linguistic, and computational issues: 8th International Conference on Conceptual Structures, ICCS 2000, Darmstadt, Germany, August 14– 18, 2000: Proceedings, volume 1867 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67859-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2 .I58 2000.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1867.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67859-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1867.

Gascuel:2001:ABF

[GM01] Olivier Gascuel and B. M. E. Moret, editors. Algorithms in bioinformatics: First International Workshop, WABI 2001, Aarhus, Denmark, August 28–31, 2001 proceedings, volume 2149 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42516-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R858.A2 A44 2000; QA267.A1 L43 no.2149; R858.A2 A44 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2149.htm;
                                         http://www.springerlink.

com/content/978-3-540-
42516-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2149.

George:2003:FMS

[GM03] Chris George and Huaikou Miao, editors. Formal methods and software engineering: 4th International Conference on Formal Engineering Methods, ICFEM 2002, Shanghai, China, October 21–25, 2002: Proceedings, volume 2495 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00029-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 I5788 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2495.
htm;   http://www.springerlink.
com/content/978-3-540-
00029-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2495.

Ghosh:2004:DCI

[GM04a] R. K. Ghosh and Hrushikesha Mohanty, editors. Distributed computing and Internet technology: first international conference, ICDCIT 2004, Bhubaneswar, India, December 22–24, 2004: Proceedings, volume 3347 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24075-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I24 2004.

URL http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3347;
 http://www.springerlink.
com/content/978-3-540-
24075-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3347;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104418.

Graf:2004:MCS

[GM04b] Susanne Graf and Laurent Mounier, editors. Model Checking Software: 11th International SPIN Workshop, Barcelona, Spain, April 1–3, 2004: Proceedings, volume 2989 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21314-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 I58 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2989.
htm;   http://www.springerlink.
com/content/978-3-540-
21314-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2989;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96721.

Gschwind:2005:SEM

[GM05] Thomas Gschwind and Cecilia Mascolo, editors. Software engineering and middleware: 4th international workshop, SEM 2004, Linz, Austria, September 20–21, 2004. Revised selected papers, volume 3437 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25328-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25328-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3437;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107130.

Goble:2006:DLA

[GM06a] Lou Goble and John-Jules Ch. Meyer, editors. Deontic Logic and Artificial Normative Systems: 8th International Workshop on Deontic Logic in Computer Science, DEON 2006, Utrecht, The Netherlands, July 12–14, 2006. Proceedings, volume 4048 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35842-0 (print), 3-540-35843-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        35843-5.

Goubin:2006:CHE

[GM06b] Louis Goubin and Mitsuru Matsui, editors. Cryptographic Hardware and Embedded Systems — CHES 2006: 8th International Workshop, Yokohama, Japan, October 10–13, 2006. Proceedings, volume 4249 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46559-6 (print), 3-540-46561-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46561-4.

Gelbukh:2007:MAA

[GM07a] Alexander Gelbukh and Ángel Fernando Kuri Morales, editors. MI-CAI 2007: Advances in Artificial Intelligence: 6th Mexican International Conference on Artificial Intelligence, Aguascalientes, Mexico, November 4–10, 2007. Proceedings, volume 4827 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76630-8 (print), 3-540-76631-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76631-5.

Gurevich:2007:TPF

[GM07b] Yuri Gurevich and Bertrand Meyer, editors. Tests and Proofs: First International Conference, TAP 2007, Zurich, Switzerland, February 12–13, 2007. Revised Papers, volume 4454 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73769-3 (print), 3-540-73770-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73770-4.

Gelbukh:2008:MAA

[GM08a] Alexander Gelbukh and Eduardo F. Morales, editors. MICAI 2008: Advances in Artificial Intelligence: 7th Mexican International Conference on Artificial Intelligence, Atizapán de Zaragoza, Mexico, October 27–31, 2008 Proceedings, volume 5317 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88635-4 (print), 3-540-88636-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88636-5.

Gupta:2008:CAV

[GM08b] Aarti Gupta and Sharad Malik, editors. Computer Aided Verification: 20th International Conference, CAV 2008 Princeton, NJ, USA, July 7–14, 2008 Proceedings, volume 5123 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70543-0 (print), 3-540-70545-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70545-1.

Gupta:2010:MAI

[GM10] Indranil Gupta and Cecilia Mascolo, editors. Middleware 2010: ACM/ IFIP/USENIX 11th International Middleware Conference, Bangalore, India, November 29 — December 3, 2010. Proceedings, volume 6452 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16954-6 (print), 3-642-16955-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16955-7.

Giancarlo:2011:CPM

[GM11a] Raffaele Giancarlo and Giovanni Manzini, editors. Combinatorial Pattern Matching: 22nd Annual Symposium, CPM 2011, Palermo, Italy, June 27–29, 2011. Proceedings, volume 6661 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21457-6 (print), 3-642-21458-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21458-5.

Groce:2011:MCS

[GM11b] Alex Groce and Madanlal Musuvathi, editors. Model Checking Software: 18th International SPIN Workshop, Snowbird, UT, USA, July 14–15, 2011. Proceedings, volume 6823 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22305-2 (print), 3-642-22306-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22306-8.

Giannakopoulou:2012:FFM

[GM12] Dimitra Giannakopoulou and Dominique Méry, editors. FM 2012: Formal Methods: 18th International Symposium, Paris, France, August 27–31, 2012. Proceedings, volume 7436 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32758-3 (print), 3-642-32759-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32759-9.

Gil:2005:SWI

[GMBM05] Yolanda Gil, Enrico Motta, V. Richard Benjamins, and Mark A. Musen, editors. The Semantic Web — ISWC 2005: 4th International Semantic Web Conference, ISWC 2005, Galway, Ireland, November 6–10, 2005, Proceedings, volume 3729 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29754-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29754-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3729.

Gobel:2006:TID

[GMI06] Stefan Göbel, Rainer Malkewitz, and Ido Iurgel, editors. Technologies for Interactive Digital Storytelling and Entertainment: Third International Conference, TIDSE 2006, Darmstadt, Germany, December 4–6, 2006. Proceedings, volume 4326 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49934-2 (print), 3-540-49935-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49935-0.

Gao:2008:MII

[GML^{+}08a] Xiaohong Gao, Henning Müller, Martin J. Loomes, Richard Comley, and Shuqian Luo, editors. Medical Imaging and Informatics: 2nd International Conference, MIMI 2007, Beijing, China, August 14–16, 2007 Revised Selected Papers, volume 4987 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79489-1 (print), 3-540-79490-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79490-5.

Gervasi:2008:CSAa

[GML^{+}08b] Osvaldo Gervasi, Beniamino Murgante, Antonio Laganà, David Taniar, Youngsong Mun, and Marina L. Gavrilova, editors. Computational Science and Its Applications — ICCSA 2008: International Conference, Perugia, Italy, June 30 — July 3, 2008, Proceedings, Part I, volume 5072 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69838-8 (print), 3-540-69839-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69839-5.

Gervasi:2008:CSAb

[GML^{+}08c] Osvaldo Gervasi, Beniamino Murgante, Antonio Laganà, David Taniar, Youngsong Mun, and Marina L. Gavrilova, editors. Computational Science and Its Applications — ICCSA 2008: International Conference, Perugia, Italy, June 30– July 3, 2008, Proceedings, Part II, volume 5073 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69840-X (print), 3-540-69848-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69848-7.

Giorgini:2003:AOS

[GMO03] Paolo Giorgini, Jörg P. Müller, and James Odell, editors. Agent-Oriented Software Engineering IV: 4th International Workshop, AOSE 2003, Melbourne, Australia, July 15, 2003: Revised Papers, volume 2935 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20826-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .A57 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2935.htm;
http://www.springerlink.
com/content/978-3-540-
20826-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2935;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95187.

Garay:2009:CNS

[GMO09] Juan A. Garay, Atsuko Miyaji, and Akira Otsuka, editors. Cryptology and Network Security: 8th International Conference, CANS 2009, Kanazawa, Japan, December 12–14, 2009. Proceedings, volume 5888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10432-0 (print), 3-642-10433-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10433-6.

Gesu:2006:FLA

[GMP06] Vito Gesù, Francesco Masulli, and Alfredo Petrosino, editors. Fuzzy Logic and Applications: 5th International Workshop, WILF 2003, Naples, Italy, October 9–11, 2003. Revised Selected Papers, volume 2955 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31019-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31019-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2955.

Garcia:2003:VCP

[GMS03] Narciso García, José Martínez, and Luis Salgado, editors. Visual Content Processing and Representation: 8th International Workshop, VLBV 2003, Madrid, Spain, September 18– 19, 2003: Proceedings, volume 2849 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20081-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2849.htm;
http://www.springerlink.
com/content/978-3-540-
20081-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2849;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13938.

Gollmann:2006:CSE

[GMS06] Dieter Gollmann, Jan Meier, and Andrei Sabelfeld, editors. Computer Security — ESORICS 2006: 11th European Symposium on Research in Computer Security, Hamburg, Germany, September 18–20, 2006. Proceedings, volume 4189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44601-X (print), 3-540-44605-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
44605-7.

Gramlich:2012:ARI

[GMS12] Bernhard Gramlich, Dale Miller, and Uli Sattler, editors. Automated Reasoning: 6th International Joint Conference, IJ-CAR 2012, Manchester, UK, June 26–29, 2012. Proceedings, volume 7364 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31364-7 (print), 3-642-31365-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31365-3.

Gobel:2012:LGT

[GMUW12] Stefan Göbel, Wolfgang Müller, Bodo Urban, and Josef Wiemeyer, editors. E-Learning and Games for Training, Education, Health and Sports: 7th International Conference, Edutainment 2012 and 3rd International Conference, GameDays 2012, Darmstadt, Germany, September 18– 20, 2012. Proceedings, volume 7516 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33465-2 (print), 3-642-33466-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
33466-5.

Gee:2003:BIR

[GMV03] James C. Gee, J. B. Antoine Maintz, and Michael W. Vannier, editors. Biomedical Image Registration: Second International Workshop, WBIR 2003, Philadelphia, PA, USA, June 2003: Revised Papers, volume 2717 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20343-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R857.O6 B563 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2717.
htm;   http://www.springerlink.
com/content/978-3-540-
20343-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2717;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11804.

Ganzha:2005:CAS

[GMV05] Victor G. Ganzha, Ernst W.

Mayr, and Evgenii V. Vorozhtsov, editors. Computer algebra in scientific computing: 8th international workshop, CASC 2005, Kalamata, Greece, September 12–16, 2005: proceedings, volume 3718 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28966-6 (print), 3-540-32070-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28966-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3718.

Ganzha:2006:CAS

[GMV06] Victor G. Ganzha, Ernst W. Mayr, and Evgenii V. Vorozhtsov, editors. Computer Algebra in Scientific Computing: 9th International Workshop, CASC 2006, Chişinău, Moldova, September 11– 15, 2006. Proceedings, volume 4194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45182-X (print), 3-540-45195-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45195-2.

Ganzha:2007:CAS

[GMV07] Victor G. Ganzha, Ernst W.

Mayr, and Evgenii V. Vorozhtsov, editors. Computer Algebra in Scientific Computing: 10th International Workshop, CASC 2007, Bonn, Germany, September 16– 20, 2007. Proceedings, volume 4770 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75186-6 (print), 3-540-75187-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75187-8.

Gerdt:2009:CAS

[GMV09a] Vladimir P. Gerdt, Ernst W.

Mayr, and Evgenii V. Vorozhtsov, editors. Computer Algebra in Scientific Computing: 11th International Workshop, CASC 2009, Kobe, Japan, September 13–17, 2009. Proceedings, volume 5743 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04102-7 (print), 3-642-04103-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04103-7.

Grundspenkis:2009:ADI

[GMV09b] Janis Grundspenkis, Tadeusz Morzy, and Gottfried Vossen, editors. Advances in Databases and Information Systems: 13th East European Conference, ADBIS 2009, Riga, Latvia, September 7–10, 2009. Proceedings, volume 5739 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03972-3 (print), 3-642-03973-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03973-7.

Gudmundsson:2012:CCA

[GMV12] Joachim Gudmundsson, Julián Mestre, and Taso Viglas, editors. Computing and Combinatorics: 18th Annual International Conference, COCOON 2012, Sydney, Australia, August 20–22, 2012. Proceedings, volume 7434 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32240-9 (print), 3-642-32241-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32241-9.

Gordon:1979:ELM

[GMW79] Michael J. C. Gordon, R. Milner, and Christopher P. Wadsworth. Edinburgh LCF: a mechanised logic of computation, volume 78 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CO-DEN LNCSD9. ISBN 0-387-09724-4. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 159 pp. LCCN QA9.59 .G67. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0078.htm;
http://www.springerlink.
com/content/978-0-387-
09724-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=78.

Gjessing:1988:EEC

[GN88] S. Gjessing and Kristen Nygaard, editors. ECOOP ’88, European Conference on Object-Oriented Programming, Oslo, Norway, August 15–17, 1988, proceedings, volume 322 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 3-540-50053-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E9731 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0322.htm;
 http://www.springerlink.
com/content/978-3-540-
50053-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=322.

Gottlob:1990:ESE

[GN90] G. Gottlob and W. Nejdl, editors. Expert systems in engineering: principles and applications: International Workshop, Vienna, Austria, September 24–26, 1990, proceedings, volume 462 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-53104-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .E974 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0462.htm;
 http://www.springerlink.
com/content/978-0-387-
53104-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=462.

“Papers … presented at the International Workshop on Expert Systems in Engineering”–Pref.

Grundy:1998:TPH

[GN98] J. (Jim) Grundy and Malcolm Charles Newey, editors. Theorem proving in higher order logics: 11th International Conference, TPHOLs ’98, Canberra, Australia, September 27–October 2, 1997: Proceedings, volume 1479 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64987-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96T655 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1479.htm;
 http://www.springerlink.
com/content/978-3-540-
64987-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1479.

Grabowski:2005:FAS

[GN05] Jens Grabowski and Brian Nielsen, editors. Formal approaches to software testing: 4th International Workshop, FATES 2004, Linz, Austria, September 21, 2004: Revised selected papers, volume 3395 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25109-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 F522 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25109-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3395;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106767.

Grohe:2008:PEC

[GN08] Martin Grohe and Rolf Niedermeier, editors. Parameterized and Exact Computation: Third International Workshop, IWPEC 2008, Victoria, Canada, May 14– 16, 2008. Proceedings, volume 5018 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79722-X (print), 3-540-79723-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79723-4.

Galbraith:2012:PCI

[GN12a] Steven Galbraith and Mridul Nandi, editors. Progress in Cryptology — INDOCRYPT 2012: 13th International Conference on Cryptology in India, Kolkata, India, December 9–12, 2012. Proceedings, volume 7668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34930-7 (print), 3-642-34931-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34931-7.

Gheyi:2012:FMF

[GN12b] Rohit Gheyi and David Naumann, editors. Formal Methods: Foundations and Applications: 15th Brazilian Symposium, SBMF 2012, Natal, Brazil, September 23–28, 2012. Proceedings, volume 7498 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33295-1 (print), 3-642-33296-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33296-8.

Gibbons:2012:MPC

[GN12c] Jeremy Gibbons and Pablo Nogueira, editors. Mathematics of Program Construction: 11th International Conference, MPC 2012, Madrid, Spain, June 25–27, 2012. Proceedings, volume 7342 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31112-1 (print), 3-642-31113-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31113-0.

Guerraoui:1994:OBD

[GNR94] Rachid Guerraoui, Oscar Marius Nierstrasz, and Michel Riveill, editors. Object-based distributed programming: ECOOP ’93, workshop, Kaiserslautern, Germany, July 26–27, 1993: proceedings, volume 791 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-57932-X (USA), 3-540-57932-X (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E27 1993. DM58.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0791.htm;
 http://www.springerlink.
com/content/978-0-387-
57932-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=791.

Gaudin:2007:SDD

[GNR07] Emmanuel Gaudin, Elie Najm, and Rick Reed, editors. SDL 2007: Design for Dependable Systems: 13th International SDL Forum Paris, France, September 18– 21, 2007 Proceedings, volume 4745 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74983-7 (print), 3-540-74984-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74984-4.

Grigoras:2002:AET

[GNTF02] Dan Grigoras, Alex Nicolau, Bernard Toursel, and Bertil Folliot, editors. Advanced Environments, Tools, and Applications for Cluster Computing: NATO Advanced Research Workshop, IWCC 2001, Mangalia, Romania, September 1–6, 2001. Revised Papers, volume 2326 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43672-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 N38 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2326.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43672-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2326.

Gabbay:1994:TLF

[GO94] Dov M. Gabbay and Hans Jurgen Ohlbach, editors. Temporal logic: first international conference, ICTL ’94, Bonn, Germany, July 11–14, 1994: proceedings, volume 827 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58241-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .T4414 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0827.htm;
 http://www.springerlink.
com/content/978-3-540-
58241-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=827.

Gabbay:1996:PRI

[GO96] Dov M. Gabbay and Hans Jurgen Ohlbach, editors. Practical reasoning: International Conference on Formal and Applied Practical Reasoning, FAPR ’96, Bonn, Germany, June 3–7, 1996: proceedings, volume 1085 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61313-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I575 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1085.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61313-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1085.

Gruhn:2006:SAT

[GO06] Volker Gruhn and Flavio Oquendo, editors. Software Architecture: Third European Workshop, EWSA 2006, Nantes, France, September 4–5, 2006, Revised Selected Papers, volume 4344 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69271-1 (print), 3-540-69272-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69272-0.

Gibbons:2009:TFM

[GO09] Jeremy Gibbons and José Nuno Oliveira, editors. Teaching Formal Methods: Second International Conference, TFM 2009, Eindhoven, The Netherlands, November 2–6, 2009. Proceedings, volume 5846 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04911-7 (print), 3-642-04912-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04912-5.

Giannakopoulou:2011:FAS

[GO11] Dimitra Giannakopoulou and Fernando Orejas, editors. Fundamental Approaches to Software Engineering: 14th International Conference, FASE 2011, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2011, Saarbrücken, Germany, March 26–April 3, 2011. Proceedings, volume 6603 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19810-4 (print), 3-642-19811-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        19811-3.

Godefroid:1996:POM

[God96] Patrice Godefroid. Partialorder methods for the verification of concurrent systems: an approach to the state-explosion problem, volume 1032 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60761-7 (Berlin softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). 142 pp. LCCN QA76.58 .G63 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1032.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60761-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1032.

Godefroid:2005:MCS

[God05a] Patrice Godefroid, editor. Model checking software: 12th International SPIN Workshop, San Francisco, CA, USA, August 22–24, 2005: proceedings, volume 3639 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28195-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 I58 2005.

URL http://www.springerlink.
com/content/978-3-540-
28195-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3639.

Godo:2005:SQA

[God05b] Lluís Godo, editor. Symbolic and quantitative approaches to reasoning with uncertainty: 8th European conference, ECSQARU 2005, Barcelona, Spain, July 6–8, 2005: proceedings, volume 3571 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27326-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
27326-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3571.

Goettl:1998:ITS

[Goe98] Barry P. Goettl, editor.

Intelligent tutoring systems: 4th International Conference, ITS’98, San Antonio, Texas, USA, August 16–19, 1998: Proceedings, volume 1452 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64770-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.5 .I557 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1452.htm;
 http://www.springerlink.
com/content/978-3-540-
64770-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1452.

Gogolla:1994:EER

[Gog94] Martin Gogolla. An extended entity-relationship model: fundamentals and pragmatics, volume 767 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57648-7 (New York), 3-540-57648-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 136 pp. LCCN QA76.9.D3 G63 1994. DM39.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0767.htm;
 http://www.springerlink.

com/content/978-0-387-
57648-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=767;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b14063.

Goldblatt:1982:ALC

[Gol82] Robert Goldblatt. Axiomatising the logic of computer programming, volume 130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CODEN LNCSD9. ISBN 0-387-11210-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 304 pp. LCCN QA76.7 .G65 1982. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0130.htm;
http://www.springerlink.
com/content/978-0-387-
11210-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=130.

Goldwasser:1990:ACC

[Gol90] S. Goldwasser, editor.

Advances in cryptology

— CRYPTO ’88: proceedings, volume 403 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-97196-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0403.htm;
 http://www.springerlink.
com/content/978-0-387-
97196-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=403.

Gollmann:1994:CSE

[Gol94] Dieter Gollmann, editor. Computer security: ES-ORICS 94: Third European Symposium on Research in Computer Security, Brighton, United Kingdom, November 7– 9, 1994: proceedings, volume 875 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-58618-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E98 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0875.
htm;   http://www.springerlink.
com/content/978-0-387-
58618-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=875.

Gollmann:1996:FSE

[Gol96] Dieter Gollmann, editor. Fast software encryption: third International Workshop Cambridge, UK, February 21–23, 1996: proceedings, volume 1039 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60865-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F38 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1039.htm;
http://www.springerlink.
com/content/978-3-540-
60865-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1039.

Goldreich:2011:PTC

[Gol11a] Oded Goldreich, editor.

Property Testing: Current Research and Surveys, volume 6390 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16366-1 (print), 3-642-16367-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16367-8.

Goldreich:2011:SCC

[Gol11b] Oded Goldreich, editor. Studies in Complexity and Cryptography. Miscellanea on the Interplay between Randomness and Computation: In Collaboration with Lidor Avigad, Mihir Bellare, Zvika Brakerski, Shafi Goldwasser, Shai Halevi, Tali Kaufman, Leonid Levin, Noam Nisan, Dana Ron, Madhu Sudan, Luca Trevisan, Salil Vadhan, Avi Wigderson, David Zuckerman, volume 6650 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22669-8 (print), 3-642-22670-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22670-0.

Gomi:2001:ERI

[Gom01] Takashi Gomi, editor. Evolutionary robotics: from intelligent robotics to artificial life: International Symposium, ER 2001, Tokyo, Japan, October 18–19, 2001 proceedings, volume 2217 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42737-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ211.37 .I581 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2217.htm;
 http://www.springerlink.
com/content/978-3-540-
42737-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2217.

Papers presented at the 8th International Symposium on Evolutionary Robotics (ER 2001).

Goos:1983:DIL

[Goo83] Gerhard Goos. DIANA:

an intermediate language for ADA, volume 161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, revised edition, 1983. CODEN LNCSD9. ISBN 0-387-12695-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 201 pp. LCCN QA76.73.A16 D5 1983.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t0161.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        12695-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=161.

Goronzy:2002:RAN

[Gor02]  Silke Goronzy.  Ro-

bust adaptation to nonnative accents in automatic speech recognition, volume 2560 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00325-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 144 pp. LCCN TK7882.S65 G68 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2560.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        00325-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2560.

Also available via the

World Wide Web.

Gordon:2003:FSS

[Gor03] Andrew D. Gordon, editor. Foundations of software science and computational structures: 6th International Conference, FOSSACS 2003, held as part of the joint European Conferences on Theory and Practice of Software, ETAPS 2003, Warsaw, Poland, April 7–11, 2003: Proceedings, volume 2620 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00897-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F67 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2620.htm;
 http://www.springerlink.
com/content/978-3-540-
00897-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2620.

Also available via the World Wide Web.

Gorski:2006:CSR

[Gór06] Janusz Górski, editor.

Computer Safety, Reliability, and Security: 25th International Conference, SAFECOMP 2006, Gdansk, Poland, September 27–29, 2006. Proceedings, volume 4166 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45762-3 (print), 3-540-45763-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        45763-3.

Gordon:2010:PLS

[Gor10] Andrew D. Gordon, editor. Programming Languages and Systems: 19th European Symposium on Programming, ESOP 2010, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2010, Paphos, Cyprus, March 20–28, 2010. Proceedings, volume 6012 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11956-5 (print), 3-642-11957-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11957-6.

Goto:1983:RSS

[Got83] Eiichi Goto, editor. RIMS symposia on software science and engineering, Kyoto, 1982: proceedings, volume 147 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-11980-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .R55 1983. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0147.htm;
 http://www.springerlink.
com/content/978-0-387-
11980-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=147.

Giunchiglia:2003:AOS

[GOW03] Fausto Giunchiglia, James J.

Odell, and Gerhard Weiss, editors. Agentoriented software engineering III: Third International Workshop, AOSE 2002, Bologna, Italy, July 15, 2002: revised papers and invited contributions, volume 2585 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00713-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .A57 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2585.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00713-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2585.

Gleizes:2005:ESA

[GOZ05]  Marie-Pierre  Gleizes,

Andrea Omicini, and Franco Zambonelli, editors. Engineering Societies in the Agents World V: 5th International Workshop, ESAW 2004, Toulouse, France, October 20–22, 2004. Revised Selected and Invited Papers, volume 3451 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27330-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58. URL

                                        http://www.springerlink.
                                        com/content/978-3-540-
                                        27330-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3451.

Gross:1989:EDA

[GP89] Maurice Gross and Dominique Perrin, editors. Electronic dictionaries and automata in computational linguistics: LITP Spring School on theoretical computer science, Saint-Pierre d’Oleron, France, May 25–29, 1987: proceedings, volume 377 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51465-1 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN P98 .L331 1987. DM28.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0377.
htm;   http://www.springerlink.
com/content/978-0-387-
51465-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=377.

Goebel:1998:IDM

[GP98] Vera Goebel and Thomas Plagemann, editors. Interactive distributed multimedia systems and telecommunication services: 5th International Workshop, IDMS’98: Oslo, Norway, September 8–11, 1998: Proceedings, volume 1483 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64955-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1483. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1483.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64955-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1483.

Grosky:2002:STP

[GP02] William I. Grosky and František Plášil, editors. SOFSEM 2002: Theory and Practice of Informatics: 29th Conference on Current Trends in Theory and Practice of Informatics, Milovy, Czech Republic, November 22–29, 2002. Proceedings, volume 2540 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00145-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .S62 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2540.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00145-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2540.

Also available via the

World Wide Web.

Gagalowicz:2005:CAI

[GP05] André Gagalowicz and Wilfried Philips, editors. Computer analysis of images and patterns: 11th international conference, CAIP 2005, Versailles, France, September 5–8, 2005: proceedings, volume 3691 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28969-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .I535 2005; TA1632 .I547 2005. URL http:

//www.springerlink.
com/content/978-3-540-
28969-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3691.

Gagalowicz:2007:CVC

[GP07] André Gagalowicz and Wilfried Philips, editors. Computer Vision/ Computer Graphics Collaboration Techniques: Third International Conference, MIRAGE 2007, Rocquencourt, France, March 28–30, 2007. Proceedings, volume 4418 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71456-1 (print), 3-540-71457-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71457-6.

Galbraith:2008:PBC

[GP08a] Steven D. Galbraith and Kenneth G. Paterson, editors. Pairing-Based Cryptography — Pairing 2008: Second International Conference, Egham, UK, September 1–3, 2008. Proceedings, volume 5209 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85503-3 (print), 3-540-85538-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85538-5.

Graham:2008:ISD

[GP08b] T. C. Nicholas Graham and Philippe Palanque, editors. Interactive Systems. Design, Specification, and Verification: 15th International Workshop, DSV-IS 2008 Kingston, Canada, July 16–18, 2008 Revised Papers, volume 5136 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70568-6 (print), 3-540-70569-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70569-7.

Gagalowicz:2009:CVC

[GP09a] André Gagalowicz and Wilfried Philips, editors. Computer Vision/ Computer Graphics Collaboration Techniques: 4th International Conference, MIRAGE 2009, Rocquencourt, France, May 4–6, 2009. Proceedings, volume 5496 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01810-6 (print), 3-642-01811-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01811-4.

Godo:2009:SUM

[GP09b] Lluís Godo and Andrea Pugliese, editors. Scalable Uncertainty Management: Third International Conference, SUM 2009, Washington, DC, USA, September 28–30, 2009. Proceedings, volume 5785 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04387-9 (print), 3-642-04388-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04388-8.

Guerraoui:2009:SSS

[GP09c] Rachid Guerraoui and Franck Petit, editors. Stabilization, Safety, and Security of Distributed Systems: 11th International Symposium, SSS 2009, Lyon, France, November 3–6, 2009. Proceedings, volume 5873 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05117-0 (print), 3-642-05118-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
05118-0.

Gagalowicz:2011:CVC

[GP11] André Gagalowicz and Wilfried Philips, editors. Computer Vision/ Computer Graphics Collaboration Techniques: 5th International Conference, MIRAGE 2011, Rocquencourt, France, October 10–11, 2011. Proceedings, volume 6930 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24135-2 (print), 3-642-24136-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24136-9.

Goodloe:2012:NFM

[GP12] Alwyn E. Goodloe and Suzette Person, editors. NASA Formal Methods: 4th International Symposium, NFM 2012, Norfolk, VA, USA, April 3–5, 2012. Proceedings, volume 7226 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28890-1 (print), 3-642-28891-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28891-3.

Geffner:2008:AAI

[GPAD08] Hector Geffner, Rui Prada, Isabel Machado Alexandre, and Nuno David, editors. Advances in Artificial Intelligence — IBERAMIA 2008: 11th Ibero–American Conference on AI, Lisbon, Portugal, October 14– 17, 2008. Proceedings, volume 5290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88308-8 (print), 3-540-88309-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88309-8.

Gaiti:2006:ANF

[GPAS^{+}06] Dominique Gaïti, Guy Pujolle, Ehab Al-Shaer, Ken Calvert, Simon Dobson, Guy Leduc, and Olli Martikainen, editors. Autonomic Networking: First International IFIP TC6 Conference, AN 2006, Paris, France, September 27– 29, 2006. Proceedings, volume 4195 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45891-3 (print), 3-540-45893-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45893-7.

Gomez-Perez:2002:KEK

[GPB02] Asunción Gómez-Pérez and V. Richard Benjamins, editors. Knowledge Engineering and Knowledge Management. Ontologies and the Semantic Web: 13th International Conference, EKAW 2002, Sigüenza, Spain, October 1–4, 2002. Proceedings, volume 2473 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44268-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E36 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2473.
htm;   http://www.springerlink.
com/content/978-3-540-
44268-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2473.

Gao:2010:LCP

[GPCL10] Guang R. Gao, Lori L. Pollock, John Cavazos, and Xiaoming Li, editors. Languages and Compilers for Parallel Computing: 22nd International Workshop, LCPC 2009, Newark, DE, USA, October 8–10, 2009, Revised Selected Papers, volume 5898 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13373-8 (print), 3-642-13374-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13374-9.

Gomez-Perez:2005:SWR

[GPE05] Asunción Gómez-Pérez and Jérôme Euzenat, editors. The Semantic Web: Research and Applications: Second European Semantic Web Conference, ESWC2005 Heraklion, Crete, Greece, May 29–June 1, 2005 Proceedings, volume 3532 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26124-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26124-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3532;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136731.

Garcia-Pedrajas:2010:TAIa

[GPHF^{+}10a] Nicolás García-Pedrajas, Francisco Herrera, Colin Fyfe, José Manuel Benítez, and Moonis Ali, editors. Trends in Applied Intelligent Systems: 23rd International Conference on Industrial Engineering and Other Applications of Applied Intelligent Systems, IEA/AIE 2010, Cordoba, Spain, June 1–4, 2010, Proceedings, Part I, volume 6096 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13021-6 (print), 3-642-13022-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13022-9.

Garcia-Pedrajas:2010:TAIb

[GPHF^{+}10b] Nicolás García-Pedrajas, Francisco Herrera, Colin Fyfe, José Manuel Benítez, and Moonis Ali, editors. Trends in Applied Intelligent Systems: 23rd International Conference on Industrial Engineering and Other Applications of Applied Intelligent Systems, IEA/AIE 2010, Cordoba, Spain, June 1–4, 2010, Proceedings, Part II, volume 6097 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13024-0 (print), 3-642-13025-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13025-0.

Garcia-Pedrajas:2010:TAIc

[GPHF^{+}10c] Nicolás García-Pedrajas, Francisco Herrera, Colin Fyfe, José Manuel Benítez, and Moonis Ali, editors. Trends in Applied Intelligent Systems: 23rd International Conference on Industrial Engineering and Other Applications of Applied Intelligent Systems, IEA/AIE 2010, Cordoba, Spain, June 1–4, 2010, Proceedings, Part III, volume 6098 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13032-1 (print), 3-642-13033-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13033-5.

Guting:1999:ASD

[GPL99] Ralf Hartmut Güting, Dimitris Papadias, and Frederick H. Lochovsky, editors. Advances in spatial databases: 6th International Symposium, SSD’99: Hong Kong, China, July 20–23, 1999: Proceedings, volume 1651 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66247-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S692 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1651.htm;
 http://www.springerlink.
com/content/978-3-540-
66247-1;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1651.

Gimenez:1998:TPP

[GPM98] Eduardo Gimenez and Christine Paulin-Mohring, editors. Types for proofs and programs: International Workshop TYPES ’96, Aussois, France, December 15– 19, 1996: Selected Papers, volume 1512 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65137-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1512. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1512.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65137-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1512.

Guimaraes:2009:ABC

[GPP09] Katia S. Guimarães, Anna Panchenko, and Teresa M. Przytycka, editors. Advances in Bioinformatics and Computational Biology: 4th Brazilian Symposium on Bioinformatics, BSB 2009, Porto Alegre, Brazil, July 29–31, 2009.

Proceedings, volume 5676 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03222-2 (print), 3-642-03223-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03223-3.

Golomb:2008:STA

[GPPW08] Solomon W. Golomb, Matthew G. Parker, Alexander Pott, and Arne Winterhof, editors. Sequences and their applications — SETA 2008: 5th international conference, Lexington, KY, USA, September 14–18, 2008: proceedings, volume 5203 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. ISBN 3-540-85912-8, 3-540-85911-X. LCCN QA292 .S48 2008eb. URL http:

//www.springerlink.
com/content/978-3-540-
85912-3.

Gheorghe:2012:MCI

[GPR^{+}12] Marian Gheorghe, Gheorghe Păun, Grzegorz Rozenberg, Arto Salomaa, and Sergey Verlan, editors. Membrane Computing: 12th International Conference, CMC 2011, Fontainebleau, France, August 23–26, 2011, Revised Selected Papers, volume 7184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28023-4 (print), 3-642-28024-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28024-5.

Gorodetsky:2003:CNS

[GPS03] Vladimir I. Gorodetsky, Leonard J. Popyack, and Victor A. Skormin, editors. Computer Network Security: Second International Workshop on Mathematical Methods, Models, and Architectures for Computer Network Security, MMM-ACNS 2003, St. Petersburg, Russia, September 21–23, 2003: Proceedings, volume 2776 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40797-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I565 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2776.
htm;   http://www.springerlink.
com/content/978-3-540-
40797-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2776;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12005.

Gritzalis:2010:CSE

[GPT10] Dimitris Gritzalis, Bart Preneel, and Marianthi Theoharidou, editors. Computer Security — ESORICS 2010: 15th European Symposium on Research in Computer Security, Athens, Greece, September 20–22, 2010. Proceedings, volume 6345 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15496-4 (print), 3-642-15497-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15497-3.

Gyssens:1988:IIC

[GPV88] M. Gyssens, Jan Paredaens, and D. Van Gucht, editors. ICDT ’88: 2nd International Conference on Database Theory, Bruges, Belgium, August 31–September 2, 1988: proceedings, volume 326 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-50171-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5571 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0326.htm;
 http://www.springerlink.
com/content/978-0-387-
50171-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=326.

Gonnet:2000:LTI

[GPV00] G. H. Gonnet, Daniel Panario, and Alfredo Viola, editors. LATIN 2000: theoretical informatics: 4th Latin American symposium, Punta del Este, Uruguay, April 10–14, 2000: Proceedings, volume 1776 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67306-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1776.

URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t1776.
htm;   http://www.springerlink.
com/content/978-3-540-
67306-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1776.

Gomez-Perez:2009:SWF

[GPYD09] Asunción Gómez-Pérez, Yong Yu, and Ying Ding, editors. The Semantic Web: Fourth Asian Conference, ASWC 2009, Shanghai, China, December 6–9, 2009. Proceedings, volume 5926 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10870-9 (print), 3-642-10871-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10871-6.

Guillou:1995:ACE

[GQ95] Louis C. Guillou and J.-J. Quisquater, editors. Advances in cryptology, EUROCRYPT ’95: International Conference on the Theory and Application of Cryptographic Techniques, Saint-Malo, France, May 21–25, 1995: proceedings, volume 921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59409-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C794 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0921.
htm;   http://www.springerlink.
com/content/978-3-540-
59409-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=921.

Gopalakrishnan:2011:CAV

[GQ11] Ganesh Gopalakrishnan and Shaz Qadeer, editors. Computer Aided Verification: 23rd International Conference, CAV 2011, Snowbird, UT, USA, July 14– 20, 2011. Proceedings, volume 6806 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22109-2 (print), 3-642-22110-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22110-1.

Gottlieb:2004:ECC

[GR04] Jens Gottlieb and Günther R.

Raidl, editors. Evolutionary Computation in Combinatorial Optimization, volume 3004 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21367-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3004.htm;
http://www.springerlink.
com/content/978-3-540-
21367-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3004;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96499.

Gottlieb:2006:ECC

[GR06a] Jens Gottlieb and Günther R.

Raidl, editors. Evolutionary Computation in Combinatorial Optimization: 6th European Conference, EvoCOP 2006, Budapest, Hungary, April 10–12, 2006. Proceedings, volume 3906 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33178-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33178-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3906.

Gotzhein:2006:SAM

[GR06b] Reinhard Gotzhein and Rick Reed, editors. System Analysis and Modeling: Language Profiles: 5th International Workshop, SAM 2006, Kaiserslautern, Germany, May 31 — June 2, 2006, Revised Selected Papers, volume 4320 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68371-2 (print), 3-540-68373-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68373-5.

Giese:2012:FTD

[GR12] Holger Giese and Grigore Rosu, editors. Formal Techniques for Distributed Systems: Joint 14th IFIP WG 6.1 International Conference, FMOODS 2012 and 32nd IFIP WG 6.1 International Conference, FORTE 2012, Stockholm, Sweden, June 13– 16, 2012. Proceedings, volume 7273 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30792-2 (print), 3-642-30793-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30793-5.

Grahne:1991:PII

[Gra91] G. Grahne. The problem of incomplete information in relational databases, volume 554 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54919-6 (Berlin), 0-387-54919-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 156 pp. LCCN QA76.9.D3 G69 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0554.
htm;   http://www.springerlink.
com/content/978-3-540-
54919-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=554.

Graf:1996:TI

[Gra96] Peter Graf. Term indexing, volume 1053 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61040-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 284 pp. LCCN QA76.9.A96G73 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1053.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61040-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1053.

Gramlich:2005:FCS

[Gra05] Bernhard Gramlich, editor. Frontiers of combining systems: 5th international workshop, FROCOS 2005, Vienna Austria, September 19– 21, 2005. proceedings, volume 3717 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29051-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29051-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3717.

Georgakopoulos:2007:SOC

[GRB^{+}07] Dimitrios Georgakopoulos, Norbert Ritter, Boualem Benatallah, Christian Zirpins, George Feuerlicht, Marten Schoenherr, and Hamid R. Motahari-Nezhad, editors. Service-Oriented Computing ICSOC 2006: 4th International Conference, Chicago, IL, USA, December 4–7, 2006, Workshops Proceedings, volume 4652 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75491-1 (print), 3-540-75492-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75492-3.

Greibach:1975:TPS

[Gre75] Sheila Greibach. Theory of program structures: schemes, semantics, verification, volume 36 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CO-DEN LNCSD9. ISBN 0-387-07415-5. ISSN 0302-9743 (print), 1611-3349 (electronic). 389 pp. LCCN QA76.6 .G743. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0036.htm;
 http://www.springerlink.
com/content/978-0-387-
07415-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=36.

Greenfield:1994:DPP

[Gre94] Jonathan S. Greenfield. Distributed programming paradigms with cryptography applications, volume 870 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-58496-X. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 182 pp. LCCN QA76.9.D5 G74 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0870.htm;
 http://www.springerlink.
com/content/978-0-387-
58496-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=870.

Revision of the author’s doctoral thesis, Syracuse University.

Gelbukh:2006:MAA

[GRG06] Alexander Gelbukh and Carlos Alberto Reyes-Garcia, editors. MICAI 2006: Advances in Artificial Intelligence: 5th Mexican International Conference on Artificial Intelligence, Apizaco, Mexico, November 13– 17, 2006. Proceedings, volume 4293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49026-4 (print), 3-540-49058-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49058-6.

Goesele:2010:PRD

[GRK^{+}10] Michael Goesele, Stefan Roth, Arjan Kuijper, Bernt Schiele, and Konrad Schindler, editors. Pattern Recognition: 32nd DAGM Symposium, Darmstadt, Germany, September 22– 24, 2010. Proceedings, volume 6376 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15985-0 (print), 3-642-15986-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15986-2.

Grossman:1993:HS

[Gro93] Robert L. Grossman, editor. Hybrid systems, volume 736 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57318-6 (New York), 3-540-57318-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .H93 1993. DM86.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0736.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        57318-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=736.

Papers presented at a workshop held Oct. 19– 21, 1992 at the Technical University, Lyngby, Denmark.

Grunske:2010:CBS

[GRP10] Lars Grunske, Ralf Reussner, and Frantisek Plasil, editors. Component-Based Software Engineering: 13th International Symposium, CBSE 2010, Prague, Czech Republic, June 23–25, 2010. Proceedings, volume 6092 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13237-5 (print), 3-642-13238-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13238-4.

Guelfi:2005:SED

[GRR05] Nicolas Guelfi, Gianna Reggio, and Alexander Romanovsky, editors. Scientific Engineering of Distributed Java Applications: 4th International Workshop, FIDJI 2004, Luxembourg-Kirchberg, Luxembourg, November 24–25, 2004. Revised selected papers, volume 3409 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25053-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F53 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25053-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3409;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106620.

Goldreich:2006:TCS

[GRS06] Oded Goldreich, Arnold L. Rosenberg, and Alan L. Selman, editors. Theoretical Computer Science: Essays in Memory of Shimon Even, volume 3895 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32880-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32880-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3895.

Garijo:2002:AAI

[GRT02] Francisco J. Garijo, Jose C. Riquelme, and Miguel Toro, editors. Advances in artificial intelligence: IBERAMIA 2002: 8th Ibero–American Conference on AI, Seville, Spain, November 12– 15, 2002: Proceedings, volume 2527 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00131-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I32 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2527.
htm;   http://www.springerlink.
com/content/978-3-540-
00131-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2527.

Also available via the World Wide Web.

Gruska:1977:MFC

[Gru77] Jozef Gruska, editor.

Mathematical foundations of computer science, 1977: proceedings, 6th Symposium, Tatranska Lomnica, September 5–9, 1977, volume 53 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CO-DEN LNCSD9. ISBN 0-387-08353-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.M371. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0053.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        08353-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=53.

Grumberg:1997:CAV

[Gru97] Orna Grumberg, editor. Computer aided verification: 9th international conference, CAV’97, Haifa, Israel, June 22–25, 1997: proceedings, volume 1254 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63166-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 C38 1997.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1254.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63166-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1254.

Gruhn:1998:SPT

[Gru98] Volker Gruhn, editor. Software process technology: 6th European Workshop, EWSPT’98: Weybridge, UK, September 16–18, 1998: Proceedings, volume 1487 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64956-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1487. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1487.htm;
 http://www.springerlink.
com/content/978-3-540-
64956-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1487.

Gruska:1986:MFC

[GRW86] Jozef Gruska, B. Rovan, and J. Wiedermann, editors. Mathematical foundations of computer science 1986: proceedings of the 12th symposium, Bratislava, Czechoslovakia, August 25–29, 1986, volume 233 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16783-8 (paperback), 3-540-16783-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .M391 1986. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t0233.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        16783-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=233.

Papers presented at the 12th Symposium on Mathematical Foundations of Computer Science.

Gottler:1988:GTC

[GS88] Herbert Gottler and Hans Jurgen Schneider, editors. Graphtheoretic concepts in computer science: International Workshop WG ’87, Kloster Banz/ Staffelstein, FRG, June 29–July 1, 1987: proceedings, volume 314 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-19422-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G7171 1988. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0314.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        19422-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=314.

Based on the 13th International Workshop on Graph-Theoretic Concepts in Computer Science, organized by the Institut für Mathematische Maschinen und Datenverarbeitung of Friedrich-Alexander-Universität Erlangen-Nurnberg.

Gunther:1991:ASD

[GS91] Oliver Gunther and H.-J Schek, editors. Advances in spatial databases: 2nd Symposium, SSD ’91, Zürich, Switzerland, August 28–30, 1991: proceedings, volume 525 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54414-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S978 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0525.htm;
http://www.springerlink.
com/content/978-0-387-
54414-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=525.

Gulliver:1994:ITA

[GS94] T. Aaron Gulliver and Norman P. Secord, editors. Information theory and applications: third Canadian workshop, Rockland, Ontario, Canada, May 30– June 2, 1993: proceedings, volume 793 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57936-2 (Berlin), 0-387-57936-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q350 .C36 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0793.
htm;   http://www.springerlink.
com/content/978-3-540-
57936-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=793.

Gori:1995:TAI

[GS95] Marco Gori and Giovanni Soda, editors. Topics in artificial intelligence: Fourth Conference of the Italian Association for Artificial Intelligence, AI*IA ’95, Florence, Italy, October 11–13, 1995: proceedings, volume 992 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60437-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN Q334 .I82 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0992.
htm;   http://www.springerlink.
com/content/978-3-540-
60437-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=992.

Giancarlo:2000:CPM

[GS00a] Raffaele Giancarlo and David Sankoff, editors. Combinatorial pattern matching: 11th annual symposium, CPM 2000, Montréal, Canada, June 21–23, 2000: Proceedings, volume 1848 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67633-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1848.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1848.
htm;   http://www.springerlink.
com/content/978-3-540-
67633-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1848.

Graf:2000:TAC

[GS00b] Susanne Graf and Michael I.

Schwartzbach, editors. Tools and algorithms for the construction and analysis of systems: 6th International Conference, TACAS 2000, held as part of the Joint European Conferences on Theory and Practices of Software, ETAPS 2000, Berlin, Germany, March/April 2000: Proceedings, volume 1785 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67282-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1785. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1785.htm;
 http://www.springerlink.
com/content/978-3-540-
67282-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1785.

Gascuel:2001:CBF

[GS01] Olivier Gascuel and Marie-France Sagot, editors. Computational biology: First International Conference on Biology, Informatics, and Mathematics, JOBIM 2000 Montpellier, France, May 3–5, 2000: Selected Papers, volume 2066 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42242-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .I557 2000; QA267.A1 L43 no.2066. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2066.htm;
 http://www.springerlink.
com/content/978-3-540-
42242-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2066.

Held May 3–5, 2000.

Goethals:2005:KDI

[GS05] Bart Goethals and Arno Siebes, editors. Knowledge discovery in inductive databases: third international workshop, KDID 2004, Pisa, Italy, September 20, 2004: Revised selected and invited papers, volume 3377 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25082-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 K45 2004.

URL http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=3377;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        25082-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3377;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106731.

Guelfi:2006:RIS

[GS06] Nicolas Guelfi and Anthony Savidis, editors. Rapid Integration of Software Engineering Techniques: Second International Workshop, RISE 2005, Heraklion, Crete, Greece, September 8–9, 2005, Revised Selected Papers, volume 3943 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-34063-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34063-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3943.

Grimaud:2008:SCR

[GS08] Gilles Grimaud and François-Xavier Standaert, editors. Smart Card Research and Advanced Applications: 8th IFIP WG 8.8/11.2 International Conference, CARDIS 2008, London, UK, September 8– 11, 2008. Proceedings, volume 5189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85892-X (print), 3-540-85893-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85893-5.

Ghilardi:2009:FCS

[GS09a] Silvio Ghilardi and Roberto Sebastiani, editors. Frontiers of Combining Systems: 7th International Symposium, Fro-CoS 2009, Trento, Italy, September 16–18, 2009. Proceedings, volume 5749 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04221-X (print), 3-642-04222-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04222-5.

Gill:2009:PAD

[GS09b] Andy Gill and Terrance Swift, editors. Practical Aspects of Declarative Languages: 11th International Symposium, PADL 2009, Savannah, GA, USA, January 19– 20, 2009. Proceedings, volume 5418 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-92994-0 (print), 3-540-92995-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        92995-6.

Governatori:2010:DLC

[GS10] Guido Governatori and Giovanni Sartor, editors. Deontic Logic in Computer Science: 10th International Conference, DEON 2010, Fiesole, Italy, July 7–9, 2010. Proceedings, volume 6181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14182-X (print), 3-642-14183-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14183-6.

Graham:1978:OSA

[GSB78] Robert M. Graham, G. Seegmuller, and Rudolf Bayer. Operating systems: an advanced course, volume 60 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1978. CO-DEN LNCSD9. ISBN 0-387-08755-9. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 593 pp. LCCN QA76.6 .O63 1977. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0060.htm;
http://www.springerlink.
com/content/978-0-387-
08755-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=60.

Republished as [BSG84].

Gobel:2004:TID

[GSH^{+}04] Stefan Göbel, Ulrike Spierling, Anja Hoffmann, Ido Iurgel, Oliver Schneider, Johanna Dechau, and Axel Feix, editors. Technologies for Interactive Digital Storytelling and Entertainment: Second International Conference, TIDSE 2004, Darmstadt, Germany, June 24–26, 2004, Proceedings, volume 3105 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22283-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3105.htm;
 http://www.springerlink.
com/content/978-3-540-
22283-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3105;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98252.

Golumbic:2012:GTC

[GSLM12] Martin Charles Golumbic, Michal Stern, Avivit Levy, and Gila Morgenstern, editors. Graph-Theoretic Concepts in Computer Science: 38th International Workshop, WG 2012, Jerusalem, Israel, June 26–28, 2012, Revised Selected Papers, volume 7551 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34610-3 (print), 3-642-34611-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34611-8.

Gorodetski:2001:IAC

[GSP01] Vladimir I. Gorodetski, Victor A. Skormin, and Leonard J. Popyack, editors. Information assurance in computer networks: methods, models, and architectures for network security: International Workshop MMM-ACNS 2001, St. Petersburg, Russia, May 21– 23, 2001, volume 2052 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42103-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 M65 2001; QA267.A1 L43 no.2052. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2052.htm;
http://www.springerlink.
com/content/978-3-540-
42103-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2052.

Grieskamp:2000:IFM

[GSS00] Wolfgang Grieskamp, Thomas Santen, and Bill Stoddart, editors. Integrated Formal Methods: Second International Conference, IFM 2000, Dagstuhl Castle, Germany, November 1–3, 2000. Proceedings, volume 1945 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41196-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 I58 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1945.htm;
 http://www.springerlink.
com/content/978-3-540-
41196-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1945.

Goel:2009:DCI

[GSS09] Ashish Goel, Friedrich C. Simmel, and Petr Sosík, editors. DNA Computing: 14th International Meeting on DNA Computing, DNA 14, Prague, Czech Republic, June 2– 9, 2008. Revised Selected Papers, volume 5347 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03075-0 (print), 3-642-03076-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03076-5.

Grossi:2011:SPI

[GSS11] Roberto Grossi, Fabrizio Sebastiani, and Fabrizio Silvestri, editors. String Processing and Information Retrieval: 18th International Symposium, SPIRE 2011, Pisa, Italy, October 17–21, 2011. Proceedings, volume 7024 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24582-X (print), 3-642-24583-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24583-1.

Grumbach:2005:ACS

[GSV05]  Stéphane  Grumbach,

Liying Sui, and Victor Vianu, editors. Advances in computer science– Asian 2005: data management on the Web: 10th Asian Computing Science Conference, Kunming, China, December 7–9, 2005: proceedings, volume 3818 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30767-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .A735

                                        2005. URL http://www.
                                        loc.gov/catdir/toc/
                                        fy0604/2005937050.html;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        30767-9.

Ganter:2005:FCAb

[GSW05] Bernhard Ganter, Gerd Stumme, and Rudolf Wille, editors. Formal Concept Analysis: Foundations and Applications, volume 3626 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27891-5. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 348 pp. LCCN QA171.5 .F67 2005. URL

                                        http://springerlink.
                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3626;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        27891-7.

Grass:2006:ACS

[GSW06] Werner Grass, Bernhard Sick, and Klaus Waldschmidt, editors. Architecture of Computing Systems — ARCS 2006: 19th International Conference, Frankfurt/ Main, Germany, March 13–16, 2006. Proceedings, volume 3894 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32765-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32765-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3894.

Giles:2010:ASN

[GSYZ10] Lee Giles, Marc Smith, John Yen, and Haizheng Zhang, editors. Advances in Social Network Mining and Analysis: Second International Workshop, SNAKDD 2008, Las Vegas, NV, USA, August 24–27, 2008, volume 5498 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14928-6 (print), 3-642-14929-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14929-0.

Gorny:1987:VPS

[GT87] P. Gorny and Michael J. Tauber, editors. Visualization in programming: [selected contributions], volume 282 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18507-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5261 1986. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0282.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        18507-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=282.

Gorny:1990:VHC

[GT90] P. Gorny and Michael J. Tauber, editors. Visualization in humancomputer interaction: selected contributions, volume 439 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-52698-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I595 1988. DM42.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0439.htm;
 http://www.springerlink.
com/content/978-0-387-
52698-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=439.

Grechenig:1993:HCI

[GT93] Thomas Grechenig and Manfred Tscheligi, editors. Human computer interaction: Vienna conference, VCHCI ’93, Fin de siècle, Vienna, Austria, September 20–22, 1993: proceedings, volume 733 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57312-7 (Berlin), 0-387-57312-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 V54 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0733.htm;
 http://www.springerlink.
com/content/978-3-540-
57312-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=733.

Geist:2003:CHD

[GT03] Daniel Geist and Enrico Tronci, editors. Correct Hardware Design and Verification Methods: 12th IFIP WG 10.5 Advanced Research Working Conference, CHARME 2003, L’Aquila, Italy, October 21–24, 2003: Proceedings, volume 2860 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20363-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.75. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2860.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20363-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2860;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b93958.

Giunchiglia:2004:TAS

[GT04] Enrico Giunchiglia and Armando Tacchella, editors. Theory and Applications of Satisfiability Testing: 6th International Conference, SAT 2003, Santa Margherita Ligure, Italy, May 5–8, 2003: Selected Revised Papers, volume 2919 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20851-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.3 .S365 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2919.htm;
 http://www.springerlink.
com/content/978-3-540-
20851-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2919;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95238.

Grieser:2005:IHI

[GT05] Gunter Grieser and Yuzuru Tanaka, editors. Intuitive human interfaces for organizing and accessing intellectual assets: international workshop, Dagstuhl castle, Germany, March 1–5, 2004. Revised selected papers, volume 3359 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24465-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.U83 I58 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24465-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3359;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104697.

Gavrilova:2008:TCSa

[GT08] Marina L. Gavrilova and C. J. Kenneth Tan, editors. Transactions on Computational Science I, volume 4750 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79298-8 (print), 3-540-79299-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79299-4.

Gavrilova:2009:TCSc

[GT09a] Marina L. Gavrilova and C. J. Kenneth Tan, editors. Transactions on Computational Science III, volume 5300 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00211-0 (print), 3-642-00212-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00212-0.

Gavrilova:2009:TCSf

[GT09b] Marina L. Gavrilova and C. J. Kenneth Tan, editors. Transactions on Computational Science VI, volume 5730 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10648-X (print), 3-642-10649-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10649-1.

Gavrilova:2010:TCSa

[GT10] Marina L. Gavrilova and C. J. Kenneth Tan, editors. Transactions on Computational Science VII, volume 5890 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11388-5 (print), 3-642-11389-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11389-5.

Gavrilova:2011:TCSa

[GT11a] Marina L. Gavrilova and C. J. Kenneth Tan, editors. Transactions on Computational Science VIII, volume 6260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16235-5 (print), 3-642-16236-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16236-7.

Gavrilova:2011:TCSc

[GT11b] Marina L. Gavrilova and C. J. Kenneth Tan, editors. Transactions on Computational Science XIII, volume 6750 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22618-3 (print), 3-642-22619-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22619-9.

Gavrilova:2012:TCSb

[GT12] Marina L. Gavrilova and C. J. Kenneth Tan, editors. Transactions on Computational Science XVI, volume 7380 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32662-5 (print), 3-642-32663-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32663-9.

Gavrilova:2010:TCSb

[GTA10] Marina L. Gavrilova, C. J. Kenneth Tan, and François Anton, editors. Transactions on Computational Science IX: Special Issue on Voronoi Diagrams in Science and Engineering, volume 6290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16006-9 (print), 3-642-16007-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16007-3.

Gavrilova:2009:TCSd

[GTM09a] Marina L. Gavrilova, C. J. Kenneth Tan, and Edward David Moreno, editors. Transactions on Computational Science IV: Special Issue on Security in Computing, volume 5430 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01003-2 (print), 3-642-01004-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01004-0.

Gervasi:2009:CSAa

[GTM^{+}09b] Osvaldo Gervasi, David Taniar, Beniamino Murgante, Antonio Laganà, Youngsong Mun, and Marina L. Gavrilova, editors. Computational Science and Its Applications — ICCSA 2009: International Conference, Seoul, Korea, June 29– July 2, 2009, Proceedings, Part I, volume 5592 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02453-X (print), 3-642-02454-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02454-2.

Gervasi:2009:CSAb

[GTM^{+}09c] Osvaldo Gervasi, David Taniar, Beniamino Murgante, Antonio Laganà, Youngsong Mun, and Marina L. Gavrilova, editors. Computational Science and Its Applications — ICCSA 2009: International Conference, Seoul, Korea, June 29–July 2, 2009, Proceedings, Part II, volume 5593 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02456-4 (print), 3-642-02457-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02457-3.

Gavrilova:2010:TCSc

[GTM10a] Marina L. Gavrilova, C. J. Kenneth Tan, and Edward David Moreno, editors. Transactions on Computational Science X: Special Issue on Security in Computing, Part I, volume 6340 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17498-1 (print), 3-642-17499-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17499-5.

Gavrilova:2010:TCSd

[GTM10b] Marina L. Gavrilova, C. J. Kenneth Tan, and Edward David Moreno, editors. Transactions on Computational Science XI: Special Issue on Security in Computing, Part II, volume 6480 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17696-8 (print), 3-642-17697-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17697-5.

Gavrilova:2011:TCSd

[GTM11] Marina L. Gavrilova, C. J. Kenneth Tan, and Mir Abolfazl Mostafavi, editors. Transactions on Computational Science XIV: Special Issue on Voronoi Diagrams and Delaunay Triangulation, volume 6970 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25248-6 (print), 3-642-25249-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25249-5.

Gavrilova:2012:TCSa

[GTP12] Marina L. Gavrilova, C. J. Kenneth Tan, and Cong-Vinh Phan, editors. Transactions on Computational Science XV: Special Issue on Advances in Autonomic Computing: Formal Engineering Methods for Nature-Inspired Computing Systems, volume 7050 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28524-4 (print), 3-642-28525-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        28525-7.

Gavrilova:2011:TCSb

[GTSS11] Marina L. Gavrilova, C. J. Kenneth Tan, Alexei Sourin, and Olga Sourina, editors. Transactions on Computational Science XII: Special Issue on Cyberworlds, volume 6670 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22335-4 (print), 3-642-22336-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22336-5.

Galindo:2004:DES

[GTT04] Fernando Galindo, Makoto Takizawa, and Roland Traunmüller, editors. Database and Expert Systems Applications: 15th International Conference, DEXA 2004, Zaragoza, Spain, August 30–September 3, 2004: Proceedings, volume 3180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22936-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3180.htm;
http://www.springerlink.
com/content/978-3-540-
22936-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3180;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99664.

Gonzalo:2006:RAT

[GTVC06] Julio Gonzalo, Costantino Thanos, M. Felisa Verdejo, and Rafael C. Carrasco, editors. Research and Advanced Technology for Digital Libraries: 10th European Conference, ECDL 2006, Alicante, Spain, September 17–22, 2006. Proceedings, volume 4172 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44636-2 (print), 3-540-44638-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
44638-5.

Gradel:2002:ALI

[GTW02] Erich Grädel, Wolfgang Thomas, and Thomas Wilke, editors. Automata Logics, and Infinite Games: A Guide to Current Research, volume 2500 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00388-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 385 pp. LCCN QA76.9.M35 A85 2002. URL http:

//link.springer.de/
link/service/series/
0558/tocs/t2500.htm;
 http://www.springerlink.
com/content/978-3-540-
00388-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2500.

Gavrilova:2008:TCSb

[GTW^{+}08] Marina L. Gavrilova, C. J. Kenneth Tan, Yingxu Wang, Yiyu Yao, and Guoyin Wang, editors. Transactions on Computational Science II, volume 5150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87562-X (print), 3-540-87563-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87563-5.

Gavrilova:2009:TCSe

[GTWC09] Marina L. Gavrilova, C. J. Kenneth Tan, Yingxu Wang, and Keith C. C. Chan, editors. Transactions on Computational Science V: Special Issue on Cognitive Knowledge Representation, volume 5540 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02096-8 (print), 3-642-02097-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02097-1.

Grieser:2003:DSI

[GTY03] Gunter Grieser, Yuzuru Tanaka, and Akihiro Yamamoto, editors. Discovery Science: 6th International Conference, DS 2003, Sapporo, Japan, October 17–19, 2003: Proceedings, volume 2843 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20293-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q174. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2843.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20293-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2843;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b14292.

Galil:1995:CPM

[GU95] Zvi Galil and E. Ukkonen, editors. Combinatorial pattern matching: 6th annual symposium, CPM 95, Espoo, Finland, July 5–7, 1995: proceedings, volume 937 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60044-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 1995.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0937.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60044-2;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=937.

Gudmundsson:2008:ATS

[Gud08] Joachim Gudmundsson, editor. Algorithm Theory — SWAT 2008: 11th Scandinavian Workshop on Algorithm Theory, Gothenburg, Sweden, July 2–4, 2008. Proceedings, volume 5124 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69900-7 (print), 3-540-69903-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69903-3.

Guessarian:1981:AS

[Gue81] Irene Guessarian. Algebraic semantics, volume 99 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10284-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 158 pp. LCCN QA267.3 .G84. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0099.
htm;   http://www.springerlink.
com/content/978-0-387-

10284-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=99.

Guessarian:1990:SSC

[Gue90] Irene Guessarian, editor. Semantics of systems of concurrent processes: LITP Spring School on Theoretical Computer Science, La Roche Posay, France, April 23–27, 1990, proceedings, volume 469 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-53479-2 (Berlin), 0-387-53479-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E26 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0469.htm;
 http://www.springerlink.
com/content/978-3-540-
53479-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=469.

The School was organized jointly by the LITP (Laboratoire d’informatique théorique et programmation, Paris) and IRIT (Institut pour la recherche en informatique de Toulouse).

Guerraoui:1999:EOO

[Gue99] Rachid Guerraoui, editor. ECOOP’99: object oriented programming: 13th European Conference, Lisbon, Portugal, June 14–18, 1999: Proceedings, volume 1628 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66156-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1628.htm;
 http://www.springerlink.
com/content/978-3-540-
66156-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1628.

Guerraoui:2001:MIA

[Gue01] Rachid Guerraoui, editor. Middleware 2001: IFIP/ACM International Conference on Distributed Systems Platforms, Heidelberg, Germany, November 12–16, 2001: Proceedings, volume 2218 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42800-3 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2001; QA76.76.M54 I34 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2218.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42800-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2218.

Guerraoui:2004:DCI

[Gue04] Rachid Guerraoui, editor. Distributed Computing: 18th International Conference, DISC 2004, Amsterdam, The Netherlands, October 4–7, 2004. Proceedings, volume 3274 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23306-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D438 2004.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        23306-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3274;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b101206.

Guelfi:2005:RIS

[Gue05] Nicolas Guelfi, editor. Rapid integration of software engineering techniques: first International workshop, RISE 2004, Luxembourg-Kirchberg, Luxembourg, November 26, 2004. Revised selected papers, volume 3475 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25812-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25812-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3475;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136153.

Gunther:1988:ACE

[Gün88a] Christoph G. Günther, editor. Advances in cryptology — EUROCRYPT ’88: Workshop on the Theory and Application of Cryptographic Techniques, Davos, Switzerland, May 25–27, 1988: proceedings, volume 330 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-50251-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E9641 1988.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0330.
htm;   http://www.springerlink.
com/content/978-0-387-
50251-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=330.

Sponsored by the International Association for Cryptologic Research.

Gunther:1988:ESG

[Gun88b] Oliver Gunther. Efficient structures for geometric data management, volume 337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-50463-X (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 135 pp. LCCN QA76.9.D35 G861 1988. DM27.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0337.htm;
 http://www.springerlink.
com/content/978-0-387-
50463-6;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=337.

“Revised and extended version of a Ph. D. dissertation submitted to the Department of Electrical Engineering and Computer Sciences, University of California at Berkeley”–Pref.

Gunther:1994:MCA

[Gun94] Christoph G. Gunther, editor. Mobile communications: advanced systems components: 1994 International Zürich Seminar on Digital Communications, Zürich, Switzerland, March 8– 11, 1994: proceedings, volume 783 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-57856-0 (Berlin), 0-387-57856-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK6570.M6 I59 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0783.htm;
 http://www.springerlink.
com/content/978-3-540-
57856-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=783.

Gupta:1999:PAD

[Gup99] Gopal Gupta, editor. Practical aspects of declarative languages: First International Workshop, PADL ’99, San Antonio, Texas, USA, January 18–19, 1999: Proceedings, volume 1551 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65527-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .P32 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1551.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65527-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1551.

Gupta:2010:CCI

[Gup10] Rajiv Gupta, editor. Compiler Construction: 19th International Conference, CC 2010, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2010, Paphos, Cyprus, March 20–28, 2010. Proceedings, volume 6011 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11969-7 (print), 3-642-11970-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11970-5.

Guruswami:2004:LDE

[Gur04] Venkatesan Guruswami.

List decoding of errorcorrecting codes: winning thesis of the 2002 ACM doctoral dissertation competition, volume 3282 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-24051-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
24051-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3282;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104335.

Gutknecht:1994:PLS

[Gut94] Jürg Gutknecht, editor.

Programming languages and system architectures: international conference, Zürich, Switzerland, March 2–4, 1994: proceedings, volume 782 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57840-4 (New York), 3-540-57840-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .P79 1994. DM72.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0782.
htm;   http://www.springerlink.
com/content/978-0-387-
57840-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=782.

Gottlob:1995:DTI

[GV95] G. Gottlob and Moshe Y. Vardi, editors. Database theory — ICDT’95: 5th International Conference, Prague, Czech Republic, January 11– 13, 1995: proceedings, volume 893 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 0-387-58907-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I557 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0893.
htm;   http://www.springerlink.
com/content/978-0-387-
58907-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=893.

Grumberg:2008:YMC

[GV08] Orna Grumberg and Helmut Veith, editors. 25 Years of Model Checking: History, Achievements, Perspectives, volume 5000 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69849-3 (print), 3-540-69850-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69850-0.

Gudes:2009:DAS

[GV09] Ehud Gudes and Jaideep Vaidya, editors. Data and Applications Security XXIII: 23rd Annual IFIP WG 11.3 Working Conference, Montreal, Canada, July 12– 15, 2009. Proceedings, volume 5645 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03006-8 (print), 3-642-03007-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03007-9.

Giacobini:2012:ECM

[GVB12] Mario Giacobini, Leonardo Vanneschi, and William S. Bush, editors. Evolutionary Computation, Machine Learning and Data Mining in Bioinformatics: 10th European Conference, EvoBIO 2012, Málaga, Spain, April 11–13, 2012. Proceedings, volume 7246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29065-5 (print), 3-642-29066-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29066-4.

Garcia-Vidal:2007:WSM

[GVCA07] Jorge García-Vidal and Llorenç Cerdà-Alabern, editors. Wireless Systems and Mobility in Next Generation Internet: Third International Workshop of the EURO-NGI Network of Excellence, Sitges, Spain, June 6–9, 2006, Revised Selected Papers, volume 4396 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70968-1 (print), 3-540-70969-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70969-5.

Ganzinger:1999:LPA

[GVM99] H. Ganzinger, A. Voronkov, and David McAllester, editors. Logic for programming and automated reasoning: 6th International Conference, LPAR’99, Tbilisi, Georgia, September 6–10, 1999: Proceedings, volume 1705 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66492-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1705. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t1705.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66492-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1705.

Gaber:2010:KDS

[GVO^{+}10] Mohamed Medhat Gaber, Ranga Raju Vatsavai, Olufemi A. Omitaomu, João Gama, Nitesh V. Chawla, and Auroop R. Ganguly, editors. Knowledge Discovery from Sensor Data: Second International Workshop, Sensor-KDD 2008, Las Vegas, NV, USA, August 24–27, 2008, Revised Selected Papers, volume 5840 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12518-2 (print), 3-642-12519-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12519-5.

Gasteratos:2008:CVS

[GVT08]  Antonios  Gasteratos,

Markus Vincze, and John K. Tsotsos, editors. Computer Vision Systems: 6th International Conference, ICVS 2008 Santorini, Greece, May 12–15, 2008 Proceedings, volume 5008 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79546-4 (print), 3-540-79547-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79547-6.

Guarracino:2011:EPP

[GVT^{+}11] Mario R. Guarracino, Frédéric Vivien, Jesper Larsson Träff, Mario Cannatoro, Marco Danelutto, Anders Hast, Francesca Perla, Andreas Knüpfer, Beniamino Di Martino, and Michael Alexander, editors. Euro-Par 2010 Parallel Processing Workshops: HeteroPar, HPCC, HiBB, Core-Grid, UCHPC, HPCF, PROPER, CCPI, VHPC, Ischia, Italy, August 31– September 3, 2010, Revised Selected Papers, volume 6586 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21877-6 (print), 3-642-21878-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21878-1.

Grundy:1993:ADB

[GW93] A. F. Grundy and Michael Worboys, editors. Advances in databases: 11th British National Conference on Databases, BNCOD 11, Keele, UK, July 7– 9, 1993: proceedings, volume 696 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-56921-9 (New York), 3-540-56921-9 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B75 1993. DM58.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0696.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        56921-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=696.

Gaudel:1996:FIB

[GW96] M.-C. (Marie-Claude) Gaudel and Jim Woodcock, editors. FME ’96, industrial benefit and advances in formal methods: Third International Symposium of Formal Methods Europe cosponsored by IFIP WG 14.3, Oxford, UK, March 18–22, 1996: proceedings, volume 1051 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60973-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I593 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1051.
htm;   http://www.springerlink.
com/content/978-3-540-
60973-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1051.

Gopalakrishnan:1998:FMC

[GW98] Ganesh Gopalakrishnan and Phillip Windley, editors. Formal Methods in Computer-Aided Design: Second International Conference, FM-CAD ’98, Palo Alto, CA, USA, November 4– 6, 1998. Proceedings, volume 1522 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65191-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.65 .F53 1998.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1522.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65191-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1522.

Gutknecht:2000:MPL

[GW00] Jurg Gutknecht and Wolfgang Weck, editors. Modular programming languages: Joint Modular Languages Conference, JMLC 2000, Zürich, Switzerland, September 6–8, 2000: Proceedings, volume 1897 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67958-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1897.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67958-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1897.

Geuvers:2003:TPP

[GW03] Herman Geuvers and Freek Wiedijk, editors. Types for proofs and programs: International Workshop, TYPES 2002, Berg en Dal, the Netherlands, April 24–28, 2002: Selected Papers, volume 2646 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-14031-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T96 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2646.htm;
 http://www.springerlink.
com/content/978-3-540-
14031-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2646.

Also available via the World Wide Web.

Gorrieri:2006:FMO

[GW06a] Roberto Gorrieri and Heike Wehrheim, editors. Formal Methods for Open Object-Based Distributed Systems: 8th IFIP WG 6.1 International Conference,FMOODS 2006, Bologna, Italy, June 14– 16, 2006, Proceedings, volume 4037 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34893-X (softcover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34893-1.

Grieskamp:2006:FAS

[GW06b] Wolfgang Grieskamp and Carsten Weise, editors. Formal Approaches to Software Testing: 5th International Workshop, FATES 2005, Edinburgh, UK, July 11, 2005, Revised Selected Papers, volume 3997 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34454-3 (softcover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34454-4.

Giese:2009:ARA

[GW09] Martin Giese and Arild Waaler, editors. Automated Reasoning with Analytic Tableaux and Related Methods: 18th International Conference, TABLEAUX 2009, Oslo, Norway, July 6– 10, 2009. Proceedings, volume 5607 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02715-6 (print), 3-642-02716-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02716-1.

Gogolla:2011:TPI

[GW11a] Martin Gogolla and Burkhart Wolff, editors. Tests and Proofs: 5th International Conference, TAP 2011, Zurich, Switzerland, June 30 — July 1, 2011. Proceedings, volume 6706 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21767-2 (print), 3-642-21768-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21768-5.

Gunluk:2011:IPC

[GW11b] Oktay Günlük and Gerhard J. Woeginger, editors. Integer Programming and Combinatoral Optimization: 15th International Conference, IPCO 2011, New York, NY, USA, June 15– 17, 2011. Proceedings, volume 6655 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20806-1 (print), 3-642-20807-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20807-2.

Grossklags:2012:DGT

[GW12] Jens Grossklags and Jean Walrand, editors. Decision and Game Theory for Security: Third International Conference, GameSec 2012, Budapest, Hungary, November 5–6, 2012. Proceedings, volume 7638 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34265-5 (print), 3-642-34266-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34266-0.

Garg:2009:DCN

[GWK09] Vijay Garg, Roger Wattenhofer, and Kishore Kothapalli, editors. Distributed Computing and Networking: 10th International Conference, ICDCN 2009, Hyderabad, India, January 3–6, 2009. Proceedings, volume 5408 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-92294-6 (print), 3-540-92295-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92295-7.

Gellersen:2005:PCT

[GWS05] Hans W. Gellersen, Roy Want, and Albrecht Schmidt, editors. Pervasive computing: third international conference, pervasive 2005, Munich, Germany, May 8–13, 2005, volume 3468 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26008-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26008-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3468;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b136550.

Gao:1999:ADG

[GWY99] Xiao-Shan Gao, Dongming Wang, and Lu Yang, editors. Automated deduction in geometry: Second International Workshop, ADG ’98, Beijing, China, August 1998: Proceedings, volume 1669 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66672-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1669. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1669.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66672-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1669.

Guo:2003:PDP

[GY03] Minyi Guo and Laurence Tianruo Yang, editors. Parallel and distributed processing and applications: International Symposium, ISPA 2003, Aizu-Wakamatsu, Japan, July 2–4, 2003:

Proceedings, volume 2745. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40523-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I88 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2745.htm;
 http://www.springerlink.
com/content/978-3-540-
40523-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2745.

Gardner:2004:CCT

[GY04] Philippa Gardner and Nobuko Yoshida, editors. CONCUR 2004— Concurrency Theory: 15th International Conference, London, UK, August 31–September 3, 2004, Proceedings, volume 3170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22940-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3170.htm;
 http://www.springerlink.
com/content/978-3-540-
22940-7;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3170;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100113.

Garzon:2008:DCI

[GY08] Max H. Garzon and Hao Yan, editors. DNA Computing: 13th International Meeting on DNA Computing, DNA13, Memphis, TN, USA, June 4– 8, 2007, Revised Selected Papers, volume 4848 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77961-2 (print), 3-540-77962-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77962-9.

Gratch:2006:IVA

[GYA^{+}06] Jonathan Gratch, Michael Young, Ruth Aylett, Daniel Ballin, and Patrick Olivier, editors. Intelligent Virtual Agents: 6th International Conference, IVA 2006, Marina Del Rey, CA, USA, August 21–23, 2006. Proceedings, volume 4133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37593-7 (print), 3-540-37594-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37594-4.

Guo:2006:PDP

[GYD^{+}06] Minyi Guo, Laurence T.

Yang, Beniamino Di Martino, Hans P. Zima, Jack Dongarra, and Feilong Tang, editors. Parallel and Distributed Processing and Applications: 4th International Symposium, ISPA 2006, Sorrento, Italy, December 4–6, 2006. Proceedings, volume 4330 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68067-5 (print), 3-540-68070-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68070-3.

Gyimothy:1996:CCI

[Gyi96] Tibor Gyimothy, editor.

Compiler construction: 6th International Conference, CC’96, Linkoping, Sweden, April 24– 26, 1996: proceedings, volume 1060 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61053-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1060.htm;
 http://www.springerlink.
com/content/978-3-540-
61053-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1060.

Graf:2006:ATV

[GZ06] Susanne Graf and Wenhui Zhang, editors. Automated Technology for Verification and Analysis: 4th International Symposium, ATVA 2006, Beijing, China, October 23–26, 2006. Proceedings, volume 4218 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47237-1 (print), 3-540-47238-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
47238-4.

Goldberg:2009:AAI

[GZ09] Andrew V. Goldberg and Yunhong Zhou, editors. Algorithmic Aspects in Information and Management: 5th International Conference, AAIM 2009, San Francisco, CA, USA, June 15–17, 2009. Proceedings, volume 5564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02157-3 (print), 3-642-02158-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02158-9.

Glesner:2002:FPL

[GZR02] Manfred Glesner, Peter Zipf, and M. Renovell, editors. Fieldprogrammable logic and applications: reconfigurable computing is going mainstream: 12th International Conference, FPL 2002, Montpellier, France, September 2– 4, 2002: Proceedings, volume 2438 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44108-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.G36 I48 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2438.htm;
 http://www.springerlink.
com/content/978-3-540-
44108-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2438.

Gorodetsky:2007:AIS

[GZSC07] Vladimir Gorodetsky, Chengqi Zhang, Victor A. Skormin, and Longbing Cao, editors. Autonomous Intelligent Systems: Multi-Agents and Data Mining: Second International Workshop, AIS-ADM 2007, St. Petersburg, Russia, June 3–5, 2007. Proceedings, volume 4476 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72838-4 (print), 3-540-72839-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72839-9.

Hao:2005:CIS

[H^{+}05] Yue Hao et al., editors. Computational intelligence and security: international conference, CIS 2005, Xi’an, China, December 15–19, 2005: proceedings, volume 3801–3802 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30818-0 (vol. 1: paperback), 3-540-30819-9 (vol. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q342 .C57 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30818-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3801.

Heys:2000:SAC

[HA00] Howard Heys and Carlisle Adams, editors. Selected areas in cryptography: 6th annual International Workshop, SAC’99, Kingston, Ontario, Canada, August 9–10, 1999: Proceedings, volume 1758 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67185-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1758.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1758.
htm;   http://www.springerlink.
com/content/978-3-540-
67185-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1758.

Contents: A universal encryption standard / Helena Handschuh and Serge Vaudenay — Yarrow-160: notes on the design and analysis of the Yarrow cryptographic pseudorandom number generator / John Kelsey, Bruce Schneier, and Niels Ferguson — Elliptic curve pseudorandom sequence generators / Guang Gong, Thomas A. Berson, and Douglas R. Stinson — Adaptive-attack norm for decorrelation and super-pseudorandomness / Serge Vaudenay — Guesswork and variation distance as measures of cipher security / John O. Pliam — Modeling linear characteristics of substitutionpermutation networks / Liam Keliher, Henk Meijer, and Stafford Tavares — Strong linear dependence and unbiased distribution of non-propagative vectors / Yuliang Zheng and Xian-Mo Zhang — Security of E2 against truncated differential cryptanalysis / Shiho Moriai … [et al..] — Keyschedule cryptanalysis of DEAL / John Kelsey and Bruce Schneier — Efficient evaluation of security against generalized interpolation attack / Kazumaro Aoki — Efficient implementation of cryptosystems based on non-maximal imaginary quadratic orders / Detlef Huhnlein — Improving and extending the Lim/Lee exponentiation algorithm / Biljana Cubaleska, Andreas Rieke, and Thomas Hermann — Software optimization of decorrelation module / Fabrice Noilhan — Pseudonym systems / Anna Lysyanskaya … [et al.] — Unconditionally secure proactive secret sharing scheme with combinatorial structures / Douglas R. Stinson and R. Wei — Protecting a mobile agent’s route against collusions / Dirk Westhoff … [et al.] – Photuris: design criteria / William Allen Simpson.

Hendtlass:2002:DAA

[HA02] Tim Hendtlass and Moonis Ali, editors. Developments in Applied Artificial Intelligence: 15th International Conference on Industrial and Engineering. Applications of Artificial Intelligence and Expert Systems, IEA/AIE 2002, Cairns, Australia, June 17–20, 2002. Proceedings, volume 2358 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43781-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I556 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2358.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        43781-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2358.

Habel:1992:HRG

[Hab92] A. (Annegret) Habel. Hyperedge replacement: grammars and languages, volume 643 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 0-387-56005-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 214 pp. LCCN QA76 .H22 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0643.htm;
 http://www.springerlink.
com/content/978-0-387-
56005-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=643.

Hersch:1998:EPA

[HAB98] Roger Hersch, Jacques André, and Heather Brown, editors. Electronic publishing, artistic imaging, and digital typography: 7th International Conference on Electronic Publishing, EP ’98 held jointly with the 4th International Conference on Raster Imaging and Digital Typography, RIDT ’98, St. Malo, France, March/April 1998: Proceedings, volume 1375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64298-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z250.7.I58 1998. URL

http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t1375.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        64298-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1375.

Hand:2002:PDD

[HAB02a] D. J. Hand, Niall M. Adams, and Richard J. Bolton, editors. Pattern Detection and Discovery: ESF Exploratory Workshop, London, UK, September 16–19, 2002. Proceedings, volume 2447 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44148-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 E84 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2447.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44148-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2447.

Hochet:2002:ICD

[HAB02b] Bertrand Hochet, Antonio J. Acosta, and Manuel J. Bellido, editors. Integrated Circuit Design. Power and Timing Modeling, Optimization and Simulation: 12th International Workshop, PATMOS 2002, Seville, Spain, September 11–13, 2002. Proceedings, volume 2451 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44143-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.75 .P38 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2451.htm;
 http://www.springerlink.
com/content/978-3-540-
44143-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2451.

Also available via the World Wide Web.

Hackl:1975:PMI

[Hac75] Clemens Hackl, editor.

Programming methodology: 4th Informatik Symposium, IBM Germany, Wildbad, September 25– 27, 1974, volume 23 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07131-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I471 1974. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0023.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        07131-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=23.

“Sponsored by IBM Germany and the IBM World Trade Corporation.” English or German.

Haddawy:1994:RPU

[Had94] Peter Haddawy. Representing plans under uncertainty: a logic of time, chance, and action, volume 770 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-57697-5 (New York), 3-540-57697-5 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 129 pp. LCCN Q375 .H33 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0770.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57697-8;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=770.

Haddadi:1996:CCA

[Had96] Afsaneh Haddadi. Communication and cooperation in agent systems: a pragmatic theory, volume 1056 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61044-8. ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 148 pp. LCCN QA76.9.D5 H33 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1056.htm;
http://www.springerlink.
com/content/978-3-540-
61044-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1056.

Hadzilacos:2003:AST

[Had03] Thanasis Hadzilacos, editor. Advances in Spatial and Temporal Databases: 8th International Symposium, SSTD 2003, Santorini Island, Greece, July 24–27, 2003: Proceedings, volume 2750 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40535-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S6925 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2750.htm;
 http://www.springerlink.
com/content/978-3-540-
40535-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2750.

Haeberer:1999:AMS

[Hae99] Armando Martin Haeberer, editor. Algebraic methodology and software technology: 7th International Conference, AMAST’98, Amazonia, Brazil, January 4–8, 1999: Proceedings, volume 1548 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65462-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I5713 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1548.htm;
 http://www.springerlink.
com/content/978-3-540-
65462-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1548.

Hager:2002:SBI

[Hag02] Gregory D. Hager, editor. Sensor based intelligent robots: International Workshop, Dagstuhl Castle, Germany, October 15–20, 2000: Selected Revised Papers, volume 2238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43399-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ210.3 .I64 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2238.htm;
 http://www.springerlink.
com/content/978-3-540-
43399-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2238.

Hailpern:1982:VCP

[Hai82] Brent T. (Brent Tzion)

Hailpern. Verifying concurrent processes using temporal logic, volume 129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11205-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 208 pp. LCCN QA76.6 .H295 1982. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0129.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        11205-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=129.

Hajnicz:1996:TSF

[Haj96] Elzbieta Hajnicz. Time structures: formal description and algorithmic representation, volume 1047 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60941-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 244 pp. LCCN QA76.9.A43 H35 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1047.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60941-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1047.

Halatsis:1994:PPA

[Hal94] C. Halatsis, editor. PARLE ’94: parallel architectures and languages Europe: 6th International PARLE Conference, Athens, Greece, July 4–8, 1994: proceedings, volume 817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58184-7 (Berlin), 0-387-58184-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I564 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0817.htm;
 http://www.springerlink.
com/content/978-3-540-
58184-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=817.

Hall:1996:MTS

[Hal96]  Jane Hall.  Manage-

ment of telecommunication systems and services: modelling and implementing TMN-based multi-domain management, volume 1116 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61578-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xxi + 229 pp. LCCN HE7661 .M36 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1116.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61578-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1116.

Halldorsson:2000:ATS

[Hal00] Magnus M. Halldórsson, editor. Algorithm theory — SWAT 2000: 7th Scandinavian Workshop on Algorithm Theory, Bergen, Norway, July 5–7, 2000: Proceedings, volume 1851 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67690-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1851.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1851.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67690-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1851.

Halevi:2009:ACC

[Hal09] Shai Halevi, editor. Advances in Cryptology — CRYPTO 2009:

29th Annual International Cryptology Conference, Santa Barbara, CA, USA, August 16– 20, 2009. Proceedings, volume 5677 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03355-5 (print), 3-642-03356-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03356-8.

Hammer:1989:CCH

[Ham89] D. Hammer, editor.

Compiler compilers and high speed compilation: 2nd CCHSC Workshop, Berlin, GDR, October 10–14, 1988: proceedings, volume 371 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51364-7 (USA) (W. Ger.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 C331 1988. DM37.50. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0371.htm;
 http://www.springerlink.
com/content/978-0-387-

                                        51364-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=371.

Hammer:1991:CCT

[Ham91] D. Hammer, editor. Compiler compilers: Third International Workshop CC ’90, Schwerin, FRG, October 22–24, 1990: proceedings, volume 477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-53669-8, 0-387-53669-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I54 1990. DM37.50. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0477.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        53669-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=477.

Haridi:1995:EPP

[HAM95] Seif Haridi, Khayri Ali, and Peter Magnusson, editors. EURO-PAR ’95 parallel processing: First International EURO PAR Conference, Stockholm, Sweden, August 29–31, 1995: proceedings, volume 966 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I553 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0966.htm;
 http://www.springerlink.
com/content/978-3-540-
60247-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=966.

Hamilton:2000:AAI

[Ham00] Howard J. Hamilton, editor. Advances in artificial intelligence: 13th Biennial Conference of the Canadian Society for Computational Studies of Intelligence, AI 2000, Montréal, Quebec, Canada, May 14– 17, 2000: Proceedings, volume 1822 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67557-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1822. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1822.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67557-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1822.

Hameurlain:2008:DMG

[Ham08] Abdelkader Hameurlain, editor. Data Management in Grid and Peerto-Peer Systems: First International Conference, Globe 2008, Turin, Italy, September 3, 2008. Proceedings, volume 5187 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85175-5 (print), 3-540-85176-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85176-9.

Handler:1981:CCA

[Han81] Wolfgang Handler, editor. CONPAR 81: Conference on Analysing Problem Classes and Programming for Parallel Computing, Nurnberg, June 10–12, 1981: proceedings, volume 111 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10827-0 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .C647 1981. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0111.htm;
 http://www.springerlink.
com/content/978-0-387-
10827-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=111.

Hankin:1998:PLS

[Han98] Chris Hankin, editor.

Programming languages and systems: 7th European Symposium on Programming, ESOP ’98 held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS ’98, Lisbon, Portugal, March 28–April 4, 1998: Proceedings, volume 1381 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64302-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6.E976 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1381.
htm;   http://www.springerlink.
com/content/978-3-540-

                                        64302-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1381.

Hannebauer:2002:ADR

[Han02] Markus Hannebauer, editor. Autonomous Dynamic Reconfiguration in Multi-Agent Systems: Improving the Quality and Efficiency of Collaborative Problem Solving, volume 2427 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44312-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 H36 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2427.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44312-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2427.

Hanus:2007:PAD

[Han07] Michael Hanus, editor. Practical Aspects of Declarative Languages: 9th International Symposium, PADL 2007, Nice, France, January 14–15, 2007. Proceedings, volume 4354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69608-3 (print), 3-540-69611-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69611-7.

Hanus:2009:LBP

[Han09] Michael Hanus, editor. Logic-Based Program Synthesis and Transformation: 18th International Symposium, LOP-STR 2008, Valencia, Spain, July 17–18, 2008, Revised Selected Papers, volume 5438 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00514-4 (print), 3-642-00515-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00515-2.

Hartmann:1977:CPC

[Har77] Alfred C. Hartmann.

A Concurrent PASCAL compiler for minicomputers, volume 50 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CODEN LNCSD9. ISBN 0-387-08240-9, 3-540-08240-9. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 119 pp. LCCN QA76.6 .H3831. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0050.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        08240-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=50.

Harel:1979:FOD

[Har79] David Harel. First-order dynamic logic, volume 68 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CODEN LNCSD9. ISBN 0-387-09237-4, 3-540-09237-4. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 133 pp. LCCN QA76.6 .H34 1979. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0068.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        09237-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=68.

Rev. version of the author’s thesis, M.I.T., 1978.

Harper:2001:TCT

[Har01] Robert William Harper, editor. Types in compilation: Third International Workshop, TIC 2000, Montréal, Canada, September 21, 2000: revised Selected Papers, volume 2071 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42196-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 T52 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2071.htm;
http://www.springerlink.
com/content/978-3-540-
42196-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2071.

Harris:2007:EPC

[Har07] Don Harris, editor. Engineering Psychology and Cognitive Ergonomics: 7th International Conference, EPCE 2007, Held as Part of HCI International 2007, Beijing, China, July 22–27, 2007. Proceedings, volume 4562 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73330-2 (print), 3-540-73331-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73331-7.

Harris:2009:EPC

[Har09] Don Harris, editor. Engineering Psychology and Cognitive Ergonomics: 8th International Conference, EPCE 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009. Proceedings, volume 5639 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02727-X (print), 3-642-02728-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02728-4.

Harris:2011:EPC

[Har11a] Don Harris, editor. Engineering Psychology and Cognitive Ergonomics: 9th International Conference, EPCE 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011. Proceedings, volume 6781 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21740-0 (print), 3-642-21741-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21741-8.

Hartanto:2011:HDL

[Har11b] Ronny Hartanto, editor. A Hybrid Deliberative Layer for Robotic Agents: Fusing DL Reasoning with HTN Planning in Autonomous Robots, volume 6798 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-22579-9 (print), 3-642-22580-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 155 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22580-2.

Hasan:1996:OSQ

[Has96] Waqar Hasan. Optimization of SQL queries for parallel machines, volume 1182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-62065-6. ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 133 pp. LCCN QA76.73.S67H38 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1182.htm;
 http://www.springerlink.
com/content/978-3-540-
62065-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1182.

Hausser:1986:NPN

[Hau86] Roland R. Hausser. NEWCAT: parsing natural language using leftassociative grammar, volume 231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CODEN LNCSD9. ISBN 0-387-16781-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). 540 pp. LCCN P98 .H361 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0231.htm;
 http://www.springerlink.
com/content/978-0-387-
16781-7;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=231.

Hardy:1997:RST

[HB97] Keith Hardy and Jim Briggs, editors. Reliable software technologies — Ada-Europe ’97: 1997 Ada-Europe International Conference on Reliable Software Technologies, London, UK, June 2–6, 1997: proceedings, volume 1251 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63114-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1251.htm;
http://www.springerlink.
com/content/978-3-540-
63114-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1251.

Holz:2011:DIM

[HB11] Thorsten Holz and Herbert Bos, editors. Detection of Intrusions and Malware, and Vulnerability Assessment: 8th International Conference; DIMVA 2011, Amsterdam, The Netherlands, July 7–8, 2011. Proceedings, volume 6739 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22423-7 (print), 3-642-22424-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22424-9.

Hertzberg:2007:KAA

[HBE07] Joachim Hertzberg, Michael Beetz, and Roman Englert, editors. KI 2007: Advances in Artificial Intelligence: 30th Annual German Conference on AI, KI 2007, Osnabrück, Germany, September 10– 13, 2007. Proceedings, volume 4667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74564-5 (print), 3-540-74565-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74565-5.

Hierons:2008:FMT

[HBH08] Robert M. Hierons, Jonathan P. Bowen, and Mark Harman, editors. Formal Methods and Testing: An Outcome of the FORTEST Network, Revised Selected Papers, volume 4949 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78916-2 (print), 3-540-78917-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78917-8.

Herik:2006:CGI

[HBN06] H. Jaap Herik, Yngvi Björnsson, and Nathan S. Netanyahu, editors. Computers and Games: 4th International Conference, CG 2004, Ramat-Gan, Israel, July 5– 7, 2004. Revised Papers, volume 3846 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32488-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32488-1;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3846.

Haverkort:2000:CPE

[HBS00] Boudewijn R. Haverkort, Henrik C. Bohnenkamp, and Connie U. Smith, editors. Computer performance evaluation: modelling techniques and tools: 11th International Conference, TOOLS 2000, Schaumburg, IL, USA, March 27–31, 2000: Proceedings, volume 1786 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67260-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1786. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1786.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67260-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1786.

Hermenegildo:2005:PAD

[HC05] Manuel Hermenegildo and Daniel Cabeza, editors. Practical aspects of declarative languages: 7th international symposium, PADL 2005, Long Beach, CA, USA, January 10–11, 2005: Proceedings, volume 3350 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24362-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .P32 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3350;
 http://www.springerlink.
com/content/978-3-540-
24362-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3350;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105205.

Hua:2009:AAP

[HC09] Arrems Hua and Shih-Liang Chang, editors. Algorithms and Architectures for Parallel Processing: 9th International Conference, ICA3PP 2009, Taipei, Taiwan, June 8–11, 2009. Proceedings, volume 5574 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03094-7 (print), 3-642-03095-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03095-6.

Herrero:2011:CIS

[HC11] Álvaro Herrero and Emilio Corchado, editors. Computational Intelligence in Security for Information Systems: 4th International Conference, CISIS 2011, Held at IWANN 2011, Torremolinos-Málaga, Spain, June 8–10, 2011. Proceedings, volume 6694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21322-7 (print), 3-642-21323-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21323-6.

Hornsby:2009:SIT

[HCDL09] Kathleen Stewart Hornsby, Christophe Claramunt, Michel Denis, and Gérard Ligozat, editors. Spatial Information Theory: 9th International Conference, COSIT 2009 Aber Wrac’h, France, September 21–25, 2009 Proceedings, volume 5756 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03831-X (print), 3-642-03832-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03832-7.

Hnich:2006:RAC

[HCF06] Brahim Hnich, Mats Carlsson, and François Fages, editors. Recent Advances in Constraints: Joint ERCIM/COLOGNET International Workshop on Constraint Solving and Constraint Logic Programming, CSCLP 2005, Uppsala, Sweden, June 20–22, 2005, Revised Selected and Invited Papers, volume 3978 of Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34215-X (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34215-1.

Ho:2005:AKD

[HCL05] Tu Bao Ho, David Cheung, and Huan Li, editors. Advances in Knowledge Discovery and Data Mining: 9th Pacific-Asia Conference, PAKDD 2005 Hanoi, Vietnam, May 18–20, 2005 Proceedings, volume 3518 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26076-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26076-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3518;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136725.

Hofmeister:2006:QSA

[HCR06] Christine Hofmeister, Ivica Crnković, and Ralf Reussner, editors. Quality of Software Architectures: Second International Conference on Quality of Software Architectures, QoSA 2006, Västerås, Sweden, June 27–29, 2006 Revised Papers, volume 4214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48819-7 (print), 3-540-48820-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48820-0.

Heineman:2005:CBS

[HCS^{+}05] George T. Heineman, Ivica Crnković, Heinz W. Schmidt, Judith A. Stafford, Clemens Szyperski, and Kurt Wallnau, editors. Componentbased software engineering: 8th international symposium, CBSE 2005, St. Louis, MO, USA, May 14–15, 2005, proceedings, volume 3489 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25877-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25877-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3489;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136248.

Huang:2011:AKDa

[HCS11a] Joshua Zhexue Huang, Longbing Cao, and Jaideep Srivastava, editors. Advances in Knowledge Discovery and Data Mining: 15th Pacific-Asia Conference, PAKDD 2011, Shenzhen, China, May 24– 27, 2011, Proceedings, Part I, volume 6634 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20840-1 (print), 3-642-20841-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20841-6.

Huang:2011:AKDb

[HCS11b] Joshua Zhexue Huang, Longbing Cao, and Jaideep Srivastava, editors. Advances in Knowledge Discovery and Data Mining: 15th Pacific-Asia Conference, PAKDD 2011, Shenzhen, China, May 24– 27, 2011, Proceedings, Part II, volume 6635 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20846-0 (print), 3-642-20847-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
20847-8.

Hameurlain:2002:DES

[HCT02] Abdelkader Hameurlain, Rosine Cicchetti, and Roland Traunmüller, editors. Database and expert systems applications: 13th International Conference, DEXA 2002, Aix-en-Provence, France, September 2–6, 2002: Proceedings, volume 2453 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44126-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D35844 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2453.htm;
http://www.springerlink.
com/content/978-3-540-
44126-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2453.

Also available via the World Wide Web.

Hendrickx:2011:APA

[HDBM11] Iris Hendrickx, Sobha Lalitha Devi, António Branco, and Ruslan Mitkov, editors. Anaphora Processing and Applications: 8th Discourse Anaphora and Anaphor Resolution Colloquium, DAARC 2011, Faro, Portugal, October 6–7, 2011. Revised Selected Papers, volume 7099 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25916-2 (print), 3-642-25917-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25917-3.

Hammond:1999:IFL

[HDC99] Kevin Hammond, A. J. T. Davie, and Chris Clack, editors. Implementation of functional languages: 10th International Workshop, IFL ’98, London, UK, September 9– 11, 1998: Selected Papers, volume 1595 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66229-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .I58 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1595.htm;

http://www.springerlink.
com/content/978-3-540-
66229-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1595.

Honkela:2011:ANNa

[HDGK11a] Timo Honkela, Włodzisław Duch, Mark Girolami, and Samuel Kaski, editors. Artificial Neural Networks and Machine Learning – ICANN 2011: 21st International Conference on Artificial Neural Networks, Espoo, Finland, June 14– 17, 2011, Proceedings, Part I, volume 6791 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21734-6 (print), 3-642-21735-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21735-7.

Honkela:2011:ANNb

[HDGK11b] Timo Honkela, Włodzisław Duch, Mark Girolami, and Samuel Kaski, editors. Artificial Neural Networks and Machine Learning – ICANN 2011: 21st International Conference on Artificial Neural Networks, Espoo, Finland, June 14– 17, 2011, Proceedings, Part II, volume 6792 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21737-0 (print), 3-642-21738-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21738-8.

Hu:2012:TPM

[HdL12] Zhenjiang Hu and Juan de Lara, editors. Theory and Practice of Model Transformations: 5th International Conference, ICMT 2012, Prague, Czech Republic, May 28– 29, 2012. Proceedings, volume 7307 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30475-3 (print), 3-642-30476-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30476-7.

Hutchison:2009:APN

[HDLM09] David Hutchison, Spyros Denazis, Laurent Lefevre, and Gary J. Minden, editors. Active and Programmable Networks: IFIP TC6 7th International Working Conference, IWAN 2005, Sophia Antipolis, France, November 21– 23, 2005. Revised Papers, volume 4388 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00971-9 (print), 3-642-00972-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00972-3.

Huusko:2012:EED

[HdMKS12] Jyrki Huusko, Hermann de Meer, Sonja Klingert, and Andrey Somov, editors. Energy Efficient Data Centers: First International Workshop, E 2 DC 2012, Madrid, Spain, Mai 8, 2012, Revised Selected Papers, volume 7396 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33644-2 (print), 3-642-33645-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33645-4.

Hodges:2008:LLI

[HdQ08] Wilfrid Hodges and Ruy de Queiroz, editors. Logic, Language, Information and Computation: 15th International Workshop, WoL-LIC 2008 Edinburgh, UK, July 1–4, 2008 Proceedings, volume 5110 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69936-8 (print), 3-540-69937-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69937-8.

Hebert:1995:ORC

[Heb95] Martial Hebert, editor. Object representation in computer vision: International NSF-ARPA Workshop, New York City, NY, USA, December 5–7, 1994: proceedings, volume 994 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60477-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .I68 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0994.htm;
 http://www.springerlink.
com/content/978-3-540-
60477-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=994.

Hedin:2003:CCI

[Hed03] Gorel Hedin, editor.

Compiler construction: 12th International Conference, CC 2003 held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2003, Warsaw, Poland, April 7–11, 2003: Proceedings, volume 2622 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00904-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 C32 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2622.htm;
 http://www.springerlink.
com/content/978-3-540-
00904-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2622.

Also available via the

World Wide Web.

Heering:1994:HOA

[Hee94] J. Heering, editor. Higherorder algebra, logic, and term rewriting: first International Workshop, HOA ’93, Amsterdam The Netherlands, September 23–24, 1993: Selected Papers, volume 816 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58233-9 (Berlin), 0-387-58233-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I5785 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0816.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58233-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=816.

Heisel:2012:SSA

[Hei12] Maritta Heisel, editor. Software Service and Application Engineering: Essays Dedicated to Bernd Krämer on the Occasion of His 65th Birthday, volume 7365 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30834-1 (print), 3-642-30835-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30835-2.

Held:1991:CGP

[Hel91] Martin Held. On the computational geometry of pocket machining, volume 500 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-54103-9 (USA), 3-540-54103-9. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 178 pp. LCCN TJ1225 .H45 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0500.htm;
http://www.springerlink.
com/content/978-0-387-
54103-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=500.

Helleseth:1994:ACE

[Hel94] Tor Helleseth, editor.

Advances in cryptology, EUROCRYPT ’93: Workshop on the Theory and Application of Cryptographic Techniques, Lofthus, Norway, May 23–27, 1993: proceedings, volume 765 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57600-2 (Berlin), 0-387-57600-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 1993. DM86.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0765.
htm;   http://www.springerlink.
com/content/978-3-540-
57600-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=765.

Helmert:2008:UPT

[Hel08] Malte Helmert, editor. Understanding Planning Tasks: Domain Complexity and Heuristic Decomposition, volume 4929 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77722-9 (print), 3-540-77723-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 258 (est.) pp. LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
77723-6.

Hentenryck:2006:PAD

[Hen06] Pascal Hentenryck, editor. Practical Aspects of Declarative Languages: 8th International Symposium, PADL 2006, Charleston, SC, USA, January 9–10, 2006. Proceedings, volume 3819 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30947-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30947-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3819.

Hendren:2008:CCI

[Hen08] Laurie Hendren, editor. Compiler Construction: 17th International Conference, CC 2008, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2008, Budapest, Hungary, March 29 — April 6, 2008. Proceedings, volume 4959 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78790-9 (print), 3-540-78791-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78791-4.

Hales:2003:MAB

[HENR03] David Hales, Bruce Edmonds, Emma Norling, and Juliette Rouchier, editors. Multi-Agent-Based Simulation III: 4th International Workshop, MABS 2003, Melbourne, Australia, July 14, 2003: Revised Papers, volume 2927 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20736-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M32 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2927.
htm;   http://www.springerlink.
com/content/978-3-540-
20736-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2927;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94861.

Hertzberg:1991:EWP

[Her91] J. Hertzberg, editor.

European Workshop on Planning: EWSP ’91, Sankt Augustin, FRG, March 18–19, 1991: proceedings, volume 522 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54364-3 (Berlin), 0-387-54364-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .E993 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0522.
htm;   http://www.springerlink.
com/content/978-3-540-
54364-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=522.

Herman:1992:UPG

[Her92a] I. (Ivan) Herman. The use of projective geometry in computer graphics, volume 564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55075-5 (Berlin), 0-387-55075-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 146 pp. LCCN T385 .H47 1991. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0564.
htm;   http://www.springerlink.
com/content/978-3-540-
55075-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=564.

Outgrowth of author’s thesis (Ph. D.)–University of Leiden, Netherlands, 1990.

Herrtwich:1992:NOS

[Her92b] R. G. Herrtwich, editor. Network and operating system support for digital audio and video: second International Workshop, Heidelberg, Germany, November 18–19, 1991: proceedings, volume 614 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55639-7 (Berlin), 0-387-55639-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7881.4 .N39 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0614.htm;
 http://www.springerlink.
com/content/978-3-540-
55639-8;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=614.

Hernandez:1994:QRS

[Her94] Daniel Hernandez. Qualitative representation of spatial knowledge, volume 804 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58058-1 (Berlin), 0-387-58058-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 202 pp. LCCN Q387 .H48 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0804.htm;
http://www.springerlink.
com/content/978-3-540-
58058-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=804.

Revision of thesis (Technische Universität München, 1992).

Herlihy:2000:DCI

[Her00] Maurice Herlihy, editor. Distributed Computing: 14th International Conference, DISC 2000, Toledo, Spain, October 4–6, 2000. Proceedings, volume 1914 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41143-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1914. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1914.htm;
 http://www.springerlink.
com/content/978-3-540-
41143-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1914.

Hermanns:2002:IMC

[Her02] Holger Hermanns. Interactive Markow Chains: And The Quest for Quantified Quality, volume 2428 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44261-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 217 pp. LCCN QA274.7 .H46 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2428.
htm;   http://www.springerlink.
com/content/978-3-540-
44261-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2428.

Revised version of author’s Ph.D. Thesis, University of Erlangen, Germany.

Heudin:1998:VWF

[Heu98] J. C. Heudin, editor. Virtual worlds: First International Conference, VW ’98, Paris, France, July 1–3, 1998: Proceedings, volume 1434 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64780-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 V628 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1434.htm;
http://www.springerlink.
com/content/978-3-540-
64780-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1434.

Heudin:2000:VWS

[Heu00] J. C. Heudin, editor.

Virtual worlds: Second International Conference, VW 2000, Paris, France, July 2000: Proceedings, volume 1834 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67707-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1834. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1834.htm;
 http://www.springerlink.
com/content/978-3-540-
67707-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1834.

Heuel:2004:UPG

[Heu04] Stephan Heuel. Uncertain Projective Geometry: Statistical Reasoning for Polyhedral Object Reconstruction, volume 3008 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22029-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 205 pp. LCCN QA471 .H48 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3008.htm;
 http://www.springerlink.
com/content/978-3-540-
22029-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3008;

http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97201.

Hirtle:1997:SIT

[HF97] Stephen C. Hirtle and Andrew U. Frank, editors. Spatial information theory: a theoretical basis for GIS: International Conference COSIT’97, Laurel Highlands, Pennsylvania, USA, October 15–18, 1997: proceedings, volume 1329 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63623-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.212 .C68 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1329.htm;
http://www.springerlink.
com/content/978-3-540-
63623-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1329.

Hartenstein:1996:FPL

[HG96] Reiner Hartenstein and Manfred Glesner, editors. Field-programmable logic: smart applications, new paradigms, and compilers: 6th International Workshop on Field-Programmable Logic and Applications, FPL ’96, Darmstadt, Germany, September 23–25, 1966: proceedings, volume 1142 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61730-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7868.L6 I56 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1142.
htm;   http://www.springerlink.
com/content/978-3-540-
61730-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1142.

Herzog:1998:KAA

[HG98] O. (Otthein) Herzog and A. (Andreas) Gunter, editors. KI-98: advances in artificial intelligence: 22nd Annual German Conference on Artificial Intelligence, Bremen, Germany, September 15–17, 1998: Proceedings, volume 1504 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65080-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334.G466 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1504.
htm;   http://www.springerlink.
com/content/978-3-540-
65080-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1504.

Hartenstein:2000:FPL

[HG00] Reiner W. Hartenstein and H. Grünbacher, editors. Field-Programmable Logic and Applications. The Roadmap to Reconfigurable Computing: 10th International Conference, FPL 2000, Villach, Austria, August 27–30, 2000. Proceedings, volume 1896 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67899-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7872.L64 b I58 2000; TK7895.G36 I475 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1896.
htm;   http://www.springerlink.
com/content/978-3-540-
67899-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1896.

Halevy:2002:NGI

[HG02] Alon Halevy and Avigdor Gal, editors. Next Generation Information Technologies and Systems: 5th International Workshop, NGITS 2002, Caesarea, Israel, June 24–25, 2002. Proceedings, volume 2382 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43819-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .N498 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2382.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43819-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2382.

Huang:2012:AICa

[HGBF12] De-Shuang Huang, Yong Gan, Vitoantonio Bevilacqua, and Juan Carlos Figueroa, editors. Advanced Intelligent Computing: 7th International Conference, ICIC 2011, Zhengzhou, China, August 11–14, 2011. Revised Selected Papers, volume 6838 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24727-X (print), 3-642-24728-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24728-6.

Huang:2012:AICb

[HGGG12] De-Shuang Huang, Yong Gan, Phalguni Gupta, and M. Michael Gromiha, editors. Advanced Intelligent Computing Theories and Applications. With Aspects of Artificial Intelligence: 7th International Conference, ICIC 2011, Zhengzhou, China, August 11–14, 2011, Revised Selected Papers, volume 6839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25943-X (print), 3-642-25944-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25944-9.

Huang:2012:BIC

[HGPH12] De-Shuang Huang, Yong Gan, Prashan Premaratne, and Kyungsook Han, editors. Bio-Inspired Computing and Applications: 7th International Conference on Intelligent Computing, ICIC 2011, Zhengzhou,China, August 11–14. 2011, Revised Selected Papers, volume 6840 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24552-8 (print), 3-642-24553-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24553-4.

Hobbs:2005:DPC

[HGZ05] Michael Hobbs, Andrzej Gościński, and Wanlei Zhou, editors. Distributed and parallel computing: 6th International Conference on Algorithms and Architectures for Parallel Processing, ICA3PP, Melbourne, Australia, October 2–3, 2005: proceedings, volume 3719 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29235-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I528

2005. URL http://www.
loc.gov/catdir/toc/
fy0605/2005933043.html;
http://www.springerlink.
com/content/978-3-540-
29235-7.

Horrocks:2002:SWI

[HH02] Ian Horrocks and James A.

Hendler, editors. The Semantic Web—ISWC 2002: First International Semantic Web Conference, Sardinia, Italy, June 9–12, 2002. Proceedings, volume 2342 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43760-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88815 .I89 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2342.htm;
http://www.springerlink.
com/content/978-3-540-
43760-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2342.

Huang:2003:SSS

[HH03] Shing-Tsaan Huang and Ted Herman, editors. Self-stabilizing systems: 6th International Symposium, SSS 2003, San Francisco, CA, USA, June 24–25, 2003: Proceedings, volume 2704 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40453-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S54 S77 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2704.htm;
 http://www.springerlink.
com/content/978-3-540-
40453-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2704.

Handschuh:2004:SAC

[HH04] Helena Handschuh and M. Anwar Hasan, editors. Selected Areas in Cryptography: 11th international workshop, SAC 2004, Waterloo, Canada, August 9–10, 2004: Revised selected papers, volume 3357 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24327-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 S22 2004.

URL http://springerlink.
metapress.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3357;
http://www.springerlink.
com/content/978-3-540-
24327-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3357;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105103.

Hasan:2010:AFF

[HH10] M. Anwar Hasan and Tor Helleseth, editors. Arithmetic of Finite Fields: Third International Workshop, WAIFI 2010, Istanbul, Turkey, June 27–30, 2010. Proceedings, volume 6087 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13796-2 (print), 3-642-13797-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13797-6.

Hummel:2011:PEC

[HHG11] Karin Anna Hummel, Helmut Hlavacs, and Wilfried Gansterer, editors. Performance Evaluation of Computer and Communication Systems. Milestones and Future Challenges: IFIP WG 6.3/7.3 International Workshop, PER-FORM 2010, in Honor of Günter Haring on the Occasion of His Emeritus Celebration, Vienna, Austria, October 14–16, 2010, Revised Selected Papers, volume 6821 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25574-4 (print), 3-642-25575-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25575-5.

Handler:1986:CCA

[HHJ^{+}86] Wolfgang Handler, D. Haupt, R. Jeltsch, W. Juling, and O. Lange, editors. CONPAR 86: Conference on Algorithms and Hardware for Parallel Processing, Aachen, September 17–19, 1986: proceedings, volume 237 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16811-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .C61941 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0237.htm;
http://www.springerlink.
com/content/978-0-387-
16811-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=237.

Herskovic:2012:CTI

[HHJZ12] Valeria Herskovic, H. Ulrich Hoppe, Marc Jansen, and Jürgen Ziegler, editors. Collaboration and Technology: 18th International Conference, CRIWG 2012 Raesfeld, Germany, September 16– 19, 2012 Proceedings, volume 7493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33283-8 (print), 3-642-33284-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33284-5.

Hoffmann:2006:AKA

[HhKRT06] Achim Hoffmann, Byeong ho Kang, Debbie Richards, and Shusaku Tsumoto, editors. Advances in Knowledge Acquisition and Management: Pacific Rim Knowledge Acquisition Workshop, PKAW 2006, Guilin, China, August 7–8, 2006, Revised Selected Papers, volume 4303 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68955-9 (print), 3-540-68957-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68957-7.

Huang:2007:AICb

[HHL07a] De-Shuang Huang, Laurent Heutte, and Marco Loog, editors. Advanced Intelligent Computing Theories and Applications. With Aspects of Artificial Intelligence: Third International Conference on Intelligent Computing, ICIC 2007, Qingdao, China, August 21–24, 2007. Proceedings, volume 4682 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74201-8 (print), 3-540-74205-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
74205-0.

Huang:2007:AICa

[HHL07b] De-Shuang Huang, Laurent Heutte, and Marco Loog, editors. Advanced Intelligent Computing Theories and Applications. With Aspects of Theoretical and Methodological Issues: Third International Conference on Intelligent Computing, ICIC 2007 Qingdao, China, August 21– 24, 2007 Proceedings, volume 4681 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74170-4 (print), 3-540-74171-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74171-8.

Hertling:2008:RIR

[HHLR08] Peter Hertling, Christoph M. Hoffmann, Wolfram Luther, and Nathalie Revol, editors. Reliable Implementation of Real Number Algorithms: Theory and Practice: International Seminar Dagstuhl Castle, Germany, January 8–13, 2006 Revised Papers, volume 5045 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85520-3 (print), 3-540-85521-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85521-7.

Harju:2008:WWI

[HHLS08] Jarmo Harju, Geert Heijenk, Peter Langendörfer, and Vasilios A. Siris, editors. Wired/Wireless Internet Communications: 6th International Conference, WWIC 2008 Tampere, Finland, May 28– 30, 2008 Proceedings, volume 5031 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68805-6 (print), 3-540-68807-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68807-5.

Hanus:1997:ALP

[HHM97] Michael Hanus, J. Heering, and K. (Karl)

Meinke, editors. Algebraic and logic programming: 6th international joint conference, ALP ’97-HOA ’97, Southampton, UK. September 3– 5, 1997: proceedings, volume 1298 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63459-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63.A335 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1298.
htm;   http://www.springerlink.
com/content/978-3-540-
63459-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1298.

Hameurlain:2012:DMC

[HHMT12] Abdelkader Hameurlain, Farookh Khadeer Hussain, Franck Morvan, and A. Min Tjoa, editors. Data Management in Cloud, Grid and P2P Systems: 5th International Conference, Globe 2012, Vienna, Austria, September 5–6, 2012. Proceedings, volume 7450 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32343-X (print), 3-642-32344-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32344-7.

He:2009:LRI

[HHP09] Xiangdong He, John Horty, and Eric Pacuit, editors. Logic, Rationality, and Interaction: Second International Workshop, LORI 2009, Chongqing, China, October 8–11, 2009. Proceedings, volume 5834 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04892-7 (print), 3-642-04893-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04893-7.

Hertzberger:2001:HPC

[HHW01] Bob Hertzberger, Alfons Hoekstra, and Roy D. Williams, editors. Highperformance computing and networking: 9th International Conference, HPCN Europe, 2001, Amsterdam, The Netherlands, June 25–27, 2001: Proceedings, volume 2110 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42293-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2110; QA76.88 .I57 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2110.htm;
 http://www.springerlink.
com/content/978-3-540-
42293-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2110.

Halstead:1993:PSC

[HI93] Robert H. Halstead and T. Ito, editors. Parallel symbolic computing: languages, systems, and applications: US/ Japan workshop, Cambridge, MA, USA, October 14–17, 1992: proceedings, volume 748 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57396-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P3785 1992. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0748.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57396-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=748.

Hong:2010:FSE

[HI10] Seokhie Hong and Tetsu Iwata, editors. Fast Software Encryption: 17th International Workshop, FSE 2010, Seoul, Korea, February 7–10, 2010, Revised Selected Papers, volume 6147 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13857-8 (print), 3-642-13858-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13858-4.

Hicks:2004:MIS

[Hic04] David L. Hicks, editor. Metainformatics: International Symposium, MIS 2003, Graz, Austria, September 17– 20, 2003, Revised Papers, volume 3002 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22010-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3002.htm;
 http://www.springerlink.
com/content/978-3-540-
22010-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3002;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97912.

Higashino:2005:PDS

[Hig05] Teruo Higashino, editor. Principles of Distributed Systems: 8th International Conference, OPODIS 2004, Grenoble, France, December 15–17, 2004, Revised Selected Papers, volume 3544 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27324-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 O655 2004.

URL http://www.springerlink.
com/content/978-3-540-
27324-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3544.

Hill:1978:CSV

[Hil78] Edward Hill. A comparative study of very large data bases, volume 59 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1978. CO-DEN LNCSD9. ISBN 0-387-08653-6. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 140 pp. LCCN QA76.9.D3 H54 1978. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0059.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        08653-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=59.

A revision of the author’s thesis, George Washington University, 1977.

Hill:2006:LBP

[Hil06] Patricia M. Hill, editor. Logic Based Program Synthesis and Transformation: 15th International Symposium, LOP-STR 2005, London, UK, September 7–9, 2005, Revised Selected Papers, volume 3901 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32654-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32654-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3901.

Hirst:1981:ANL

[Hir81] Graeme Hirst. Anaphora in natural language understanding: a survey, volume 119 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10858-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 128 pp. LCCN P299.A5 H5 1981. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0119.htm;
http://www.springerlink.
com/content/978-0-387-
10858-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=119.

Hirschfeld:1997:FCF

[Hir97] Rafael Hirschfeld, editor. Financial cryptography: First International Conference, FC ’97, Anguilla, British West Indies, February 24–28, 1997: proceedings, volume 1318 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63594-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1318.htm;
 http://www.springerlink.
com/content/978-3-540-
63594-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1318.

Hirschfeld:1998:FCS

[Hir98] Rafael Hirschfeld, editor. Financial Cryptography: Second International Conference, FC ’98, Anguilla, British West Indies, February 23–25, 19987: Proceedings, volume 1465 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64951-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1465.htm;
 http://www.springerlink.
com/content/978-3-540-
64951-9;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1465.

Herrmann:2005:TMT

[HIS05] Peter Herrmann, Valérie Issarny, and Simon Shiu, editors. Trust management: third international conference, ITRUST 2005, Paris, France, May 23–26, 2005. Proceedings, volume 3477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26042-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26042-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3477;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136639.

Hunt:2000:FMC

[HJ00] Warren A. Hunt, Jr. and Steven D. Johnson, editors. Formal Methods in Computer-Aided Design: Third International Conference, FMCAD 2000, Austin, TX, USA, November 1–3, 2000. Proceedings, volume 1954 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41219-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1954; TK7874.65 .F53 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1954.
htm;   http://www.springerlink.
com/content/978-3-540-
41219-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1954.

Helleseth:2012:STA

[HJ12] Tor Helleseth and Jonathan Jedwab, editors. Sequences and Their Applications — SETA 2012: 7th International Conference, Waterloo, ON, Canada, June 4–8, 2012. Proceedings, volume 7280 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30614-4 (print), 3-642-30615-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30615-0.

Huang:2012:ICTa

[HJBF12] De-Shuang Huang, Changjun Jiang, Vitoantonio Bevilacqua, and Juan Carlos Figueroa, editors. Intelligent Computing Technology: 8th International Conference, ICIC 2012, Huangshan, China, July 25–29, 2012. Proceedings, volume 7389 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31587-9 (print), 3-642-31588-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31588-6.

Huang:2009:EICa

[HJL^{+}09a] De-Shuang Huang, Kang-Hyun Jo, Hong-Hee Lee, Hee-Jun Kang, and Vitoantonio Bevilacqua, editors. Emerging Intelligent Computing Technology and Applications: 5th International Conference on Intelligent Computing, ICIC 2009, Ulsan, South Korea, September 16–19, 2009. Proceedings, volume 5754 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04069-1 (print), 3-642-04070-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04070-2.

Huang:2009:EICb

[HJL^{+}09b] De-Shuang Huang, Kang-Hyun Jo, Hong-Hee Lee, Hee-Jun Kang, and Vitoantonio Bevilacqua, editors. Emerging Intelligent Computing Technology and Applications. With Aspects of Artificial Intelligence: 5th International Conference on Intelligent Computing, ICIC 2009 Ulsan, South Korea, September 16–19, 2009 Proceedings, volume 5755 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04019-5 (print), 3-642-04020-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04020-7.

Helmy:2006:AMM

[HJMP06] Ahmed Helmy, Brendan Jennings, Liam Murphy, and Tom Pfeifer, editors. Autonomic Management of Mobile Multimedia Services: 9th IFIP/IEEE International Conference on Management of Multimedia and Mobile Networks and Services, MMNS 2006, Dublin, Ireland, October 25– 27, 2006. Proceedings, volume 4267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47654-7 (print), 3-540-47655-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47655-9.

Hochbaum:1999:RAC

[HJRS99] Dorit S. Hochbaum, Klaus Jansen, José D. P. Rolim, and Alistair Sinclair, editors. Randomization, approximation, and combinatorial optimization: algorithms and techniques: Third International Workshop on Randomization and Approximation Techniques in Computer Science, and Second International Workshop on Approximation Algorithms for Combinatorial Optimization Problems RANDOM-APPROX ’99, Berkeley, CA, USA, August 8–11, 1999: Proceedings, volume 1671 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66329-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I648 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1671.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66329-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1671.

Hlavac:2000:STP

[HJW00] Václav Hlaváč, Keith G. Jeffery, and Jiří Wiedermann, editors. SOF-SEM 2000: Theory and Practice of Informatics: 27th Conference on Current Trends in Theory and Practice of Informatics, Milovy, Czech Republic, November 25– December 2, 2000. Proceedings, volume 1963 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41348-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1963. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1963.htm;
 http://www.springerlink.
com/content/978-3-540-
41348-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1963.

Havel:1992:MFC

[HK92] Ivan M. Havel and Vaclav Koubek, editors. Mathematical foundations of computer science 1992: 17th International Symposium, Prague, Czechoslovakia, August 24–28, 1992: proceedings, volume 629 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 0-387-55808-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 M3862 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0629.
htm;   http://www.springerlink.
com/content/978-0-387-
55808-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=629.

Proceedings of the 17th Symposium on Mathematical Foundations of Computer Science.

Haring:1994:CPE

[HK94] Gunter Haring and Gabriele Kotsis, editors. Computer performance evaluation: modelling techniques and tools: 7th international conference, Vienna, Austria, May 3–6, 1994: proceedings, volume 794 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58021-2 (Berlin), 0-387-58021-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.E94 C6496 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0794.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58021-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=794.

“The 7th International Conference on Modelling Techniques and Tools for Computer Performance Evaluation”–Pref.

Hohne:1996:VBC

[HK96] K. H. (Karl Heinz) Hohne and Ron Kikinis, editors. Visualization in biomedical computing: 4th International Conference, VBC ’96, Hamburg, Germany, September 22–25, 1996: proceedings, volume 1131 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61649-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R857.O6 I578 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1131.htm;
 http://www.springerlink.
com/content/978-3-540-
61649-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1131.

Hartenstein:1998:FPL

[HK98a] Reiner Hartenstein and Andres Keevallik, editors. Field-programmable logic and applications: from FPGAs to computing paradigm: 8th International Workshop, FPL’98: Tallinn, Estonia, August 31–September 3, 1998: Proceedings, volume 1482 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64948-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.G36 I48 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1482.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64948-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1482.

Hsu:1998:CCA

[HK98b] Wen-Lian Hsu and Ming-Yang Kao, editors. Computing and combinatorics: 4th annual International Conference, COCOON ’98, Taipei, Taiwan, August, 1998: Proceedings, volume 1449 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64824-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I543 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1449.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64824-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1449.

Henzinger:2001:ESF

[HK01] T. A. Henzinger and Christoph M. Kirsch, editors. Embedded software: First International Workshop, EM-SOFT 2001, Tahoe City, CA, USA, October 8– 10, 2001: Proceedings, volume 2211 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42673-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 E47 2001; QA267.A1 L43 no.2211. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2211.htm;
 http://www.springerlink.
com/content/978-3-540-
42673-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2211.

Hagerup:2004:ATS

[HK04] Torben Hagerup and Jyrki Katajainen, editors. Algorithm Theory— SWAT 2004: 9th Scandinavian Workshop on Algorithm Theory, Humlebæk, Denmark, July 8–10, 2004, Proceedings, volume 3111 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22339-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3111.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22339-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3111;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98413.

Hartman:2005:MDA

[HK05a] Alan Hartman and David Kreische, editors. Model Driven Architecture — Foundations and Applications: First European Conference, ECMDA-FA 2005, Nuremberg, Germany, November 7– 10, 2005. Proceedings, volume 3748 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30026-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        30026-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3748.

Ho:2005:AMIb

[HK05b] Yo-Sung Ho and Hyoung Joong Kim, editors. Advances in Multimedia Information Processing– PCM 2005, volume 3768 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30040-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30040-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3768.

Ho:2005:AMIa

[HK05c] Yo-Sung Ho and Hyoung Joong Kim, editors. Advances in multimedia information processing: PCM 2005: 6th Pacific Rim Conference on Multimedia, Jeju Island, Korea, November 13–16, 2005: proceedings, volume 3767 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30027-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I333 2002. URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        30027-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3767.

Hutchison:2007:SOS

[HK07] David Hutchison and Randy H. Katz, editors. Self-Organizing Systems: Second International Workshop, IW-SOS 2007, The Lake District, UK, September 11–13, 2007. Proceedings, volume 4725 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74916-0 (print), 3-540-74917-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74917-2.

Heng:2010:PSI

[HK10a] Swee-Huay Heng and Kaoru Kurosawa, editors. Provable Security: 4th International Conference, ProvSec 2010, Malacca, Malaysia, October 13–15, 2010. Proceedings, volume 6402 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16279-7 (print), 3-642-16280-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16280-0.

Hlineny:2010:MFC

[HK10b] Petr Hliněný and Antonín Kučera, editors. Mathematical Foundations of Computer Science 2010: 35th International Symposium, MFCS 2010, Brno, Czech Republic, August 23– 27, 2010. Proceedings, volume 6281 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15154-X (print), 3-642-15155-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15155-2.

Heyden:2011:IAS

[HK11] Anders Heyden and Fredrik Kahl, editors. Image Analysis: 17th Scandinavian Conference, SCIA 2011, Ystad, Sweden, May 2011. Proceedings, volume 6688 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21226-3 (print), 3-642-21227-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21227-7.

Harth:2012:CTW

[HK12] Andreas Harth and Nora Koch, editors. Current Trends in Web Engineering: Workshops, Doctoral Symposium, and Tutorials, Held at ICWE 2011, Paphos, Cyprus, June 20–21, 2011. Revised Selected Papers, volume 7059 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27996-1 (print), 3-642-27997-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27997-3.

Kim:2010:ACS

[hKA10] Tai hoon Kim and Hojjat Adeli, editors. Advances in Computer Science and Information Technology: AST/UCMA/ ISA/ACN 2010 Conferences, Miyazaki, Japan, June 23–25, 2010. Joint Proceedings, volume 6059 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13576-5 (print), 3-642-13577-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13577-4.

Kim:2011:FGI

[hKAS^{+}11] Tai hoon Kim, Hojjat Adeli, Dominik Slezak, Frode Eika Sandnes, Xiaofeng Song, Kyo il Chung, and Kirk P. Arnett, editors. Future Generation Information Technology: Third International Conference, FGIT 2011 in Conjunction with GDC 2011, Jeju Island, Korea, December 8–10, 2011. Proceedings, volume 7105 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-27141-3 (print), 3-642-27142-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        27142-7.

Hoffmann:1993:PDS

[HKB93] Berthold Hoffmann and B. Krieg-Bruckner. Program development by specification and transformation: the PROSPEC-TRA methodology, language family, and system, volume 680 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-56733-X. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 623 pp. LCCN QA76.76.D47 P766 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0680.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        56733-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=680.

Hand:1999:AID

[HKB99] David J. Hand, Joost N. Kok, and Michael R. Berthold, editors. Advances in intelligent data analysis: Third International Symposium, IDA-99, Amsterdam, The Netherlands, August 9– 11, 1999: Proceedings, volume 1642 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66332-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA276.A1 I28 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1642.htm;
 http://www.springerlink.
com/content/978-3-540-
66332-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1642.

Hullermeier:2010:CIK

[HKH10] Eyke Hüllermeier, Rudolf Kruse, and Frank Hoffmann, editors. Computational Intelligence for Knowledge-Based Systems Design: 13th International Conference on Information Processing and Management of Uncertainty, IPMU 2010, Dortmund, Germany, June 28 — July 2, 2010. Proceedings, volume 6178 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14048-3 (print), 3-642-14049-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        14049-5.

Kim:2010:FGI

[hKhLKŚ10] Tai hoon Kim, Young hoon Lee, Byeong-Ho Kang, and Dominik Ślezak, editors. Future Generation Information Technology: Second International Conference, FGIT 2010, Jeju Island, Korea, December 13–15, 2010. Proceedings, volume 6485 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17568-6 (print), 3-642-17569-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17569-5.

Hartmann:2008:FIK

[HKI08] Sven Hartmann and Gabriele Kern-Isberner, editors. Foundations of Information and Knowledge Systems: 5th International Symposium, FoIKS 2008, Pisa, Italy, February 11–15, 2008. Proceedings, volume 4932 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77683-4 (print), 3-540-77684-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77684-0.

Hattori:2009:NFA

[HKI^{+}09] Hiromitsu Hattori, Takahiro Kawamura, Tsuyoshi Idé, Makoto Yokoo, and Yohei Murakami, editors. New Frontiers in Artificial Intelligence: JSAI 2008 Conference and Workshops, Asahikawa, Japan, June 11–13, 2008, Revised Selected Papers, volume 5447 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00608-6 (print), 3-642-00609-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00609-8.

Harju:2007:DLT

[HKL07] Tero Harju, Juhani Karhumäki, and Arto Lepistö, editors. Developments in Language Theory: 11th International Conference, DLT 2007, Turku, Finland, July 3–6, 2007. Proceedings, volume 4588 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73207-1 (print), 3-540-73208-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73208-2.

Hirsch:2012:CST

[HKLP12] Edward A. Hirsch, Juhani Karhumäki, Arto Lepistö, and Michail Prilutskii, editors. Computer Science — Theory and Applications: 7th International Computer Science Symposium in Russia, CSR 2012, Nizhny Novgorod, Russia, July 3–7, 2012. Proceedings, volume 7353 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30641-1 (print), 3-642-30642-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30642-6.

Haton:1995:ACB

[HKM95] Jean-Paul Haton, Mark Keane, and Michel Manago, editors. Advances in case-based reasoning: Second European Workshop, EWCBR-94, Chantilly, France, November 1994: Selected Papers, volume 984 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60364-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E935 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0984.
htm;   http://www.springerlink.
com/content/978-3-540-
60364-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=984.

Huang:2007:SOC

[HKM^{+}07] Jingshan Huang, Ryszard Kowalczyk, Zakaria Maamar, David Martin, Ingo Müller, Suzette Stoutenburg, and Katia P. Sycara, editors. Service-Oriented Computing: Agents, Semantics, and Engineering: AAMAS 2007 International Workshop, SO-CASE 2007, Honolulu, HI, USA, May 14, 2007. Proceedings, volume 4504 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72618-7 (print), 3-540-72619-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72619-7.

Huzar:2011:SET

[HKM^{+}11] Zbigniew Huzar, Radek Koci, Bertrand Meyer, Bartosz Walter, and Jaroslav Zendulka, editors. Software Engineering Techniques: Third IFIP TC 2 Central and East European Conference, CEE-SET 2008, Brno, Czech Republic, October 13–15, 2008, Revised Selected Papers, volume 4980 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22385-0 (print), 3-642-22386-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22386-0.

Hromkovic:2007:SAF

[HKNW07] Juraj Hromkovič, Richard Královic, Marc Nunkesser, and Peter Widmayer, editors. Stochastic algorithms: foundations and applications: 4th international symposium, SAGA 2007, Zurich, Switzerland, September 13–14, 2007: proceedings, volume 4665 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-74870-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .S24 2007. URL http:

//www.springerlink.
com/content/978-3-540-
74870-0;   http://www.
springerlink.com/openurl.
asp?genre=issue\%26issn=
0302-9743\%26volume=
4665.

Haritsa:2008:DSA

[HKP08] Jayant R. Haritsa, Ramamohanarao Kotagiri, and Vikram Pudi, editors. Database Systems for Advanced Applications: 13th International Conference, DASFAA 2008, New Delhi, India, March 19–21, 2008. Proceedings, volume 4947 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78567-1 (print), 3-540-78568-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78568-2.

Heineman:2010:RPR

[HKP10] George T. Heineman, Jan Kofron, and Frantisek Plasil, editors. Research into Practice — Reality and Gaps: 6th International Conference on the Quality of Software Architectures, QoSA 2010, Prague, Czech Republic, June 23 — 25, 2010. Proceedings, volume 6093 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13820-9 (print), 3-642-13821-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13821-8.

Holzer:2011:DCF

[HKP11] Markus Holzer, Martin Kutrib, and Giovanni Pighizzini, editors. Descriptional Complexity of Formal Systems: 13th International Workshop, DCFS 2011, Gießen/ Limburg, Germany, July 25–27, 2011. Proceedings, volume 6808 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22599-3 (print), 3-642-22600-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22600-7.

Haindl:2007:MCS

[HKR07] Michal Haindl, Josef Kittler, and Fabio Roli, editors. Multiple Classifier Systems: 7th International Workshop, MCS 2007, Prague, Czech Republic, May 23–25, 2007. Proceedings, volume 4472 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72481-8 (print), 3-540-72523-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72523-7.

Hazas:2006:LCA

[HKS06] Mike Hazas, John Krumm, and Thomas Strang, editors. Location- and Context-Awareness: Second International Workshop, LOCA 2006, Dublin, Ireland, May 10–11, 2006, Proceedings, volume 3987 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34150-1 (softcover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34150-5.

Huet:2009:SCL

[HKS09] Gérard Huet, Amba Kulkarni, and Peter Scharf, editors. Sanskrit Computational Linguistics: First and Second International Symposia Rocquencourt, France, October 29–31, 2007 Providence, RI, USA, May 15–17, 2008 Revised Selected and Invited Papers, volume 5402 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00154-8 (print), 3-642-00155-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00155-0.

Hollmen:2012:AID

[HKT12] Jaakko Hollmén, Frank Klawonn, and Allan Tucker, editors. Advances in Intelligent Data Analysis XI: 11th International Symposium, IDA 2012, Helsinki, Finland, October 25– 27, 2012. Proceedings, volume 7619 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34155-1 (print), 3-642-34156-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34156-4.

Hromkovic:2010:TFC

[HKV10] Juraj Hromkovič, Richard Královič, and Jan Vahrenhold, editors. Teaching Fundamentals Concepts of Informatics: 4th International Conference on Informatics in Secondary Schools — Evolution and Perspectives, ISSEP 2010, Zurich, Switzerland, January 13– 15, 2010. Proceedings, volume 5941 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11375-3 (print), 3-642-11376-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11376-5.

Hameurlain:2009:TLS

[HKW09] Abdelkader Hameurlain, Josef Küng, and Roland Wagner, editors. Transactions on Large-Scale Data- and Knowledge-Centered Systems I, volume 5740 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03721-6 (print), 3-642-03722-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03722-1.

Hameurlain:2010:TLS

[HKW^{+}10] Abdelkader Hameurlain, Josef Küng, Roland Wagner, Torben Bach Pedersen, and A. Min Tjoa, editors. Transactions on Large-Scale Data- and Knowledge-Centered Systems II, volume 6380 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16174-X (print), 3-642-16175-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16175-9.

Hameurlain:2011:TLSa

[HKW11a] Abdelkader Hameurlain, Josef Küng, and Roland Wagner, editors. Transactions on Large-Scale Data- and Knowledge-Centered Systems III: Special Issue on Data and Knowledge Management in Grid and P2P Systems, volume 6790 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23073-3 (print), 3-642-23074-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23074-5.

Hameurlain:2011:TLSb

[HKW^{+}11b] Abdelkader Hameurlain, Josef Küng, Roland Wagner, Christian Böhm, Johann Eder, and Claudia Plant, editors. Transactions on Large-Scale Data- and Knowledge-Centered Systems IV: Special Issue on Database Systems for Biomedical Applications, volume 6990 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23739-8 (print), 3-642-23740-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23740-9.

Hameurlain:2012:TLSa

[HKW12a] Abdelkader Hameurlain, Josef Küng, and Roland Wagner, editors. Transactions on Large-Scale Data- and Knowledge-Centered Systems V, volume 7100 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28147-8 (print), 3-642-28148-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28148-8.

Hameurlain:2012:TLSb

[HKW^{+}12b] Abdelkader Hameurlain, Josef Küng, Roland Wagner, Stephen W. Liddle, Klaus-Dieter Schewe, and Xiaofang Zhou, editors. Transactions on Large-Scale Data- and Knowledge-Centered Systems VI: Special Issue on Database- and Expert-Systems Applications, volume 7600 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34178-0 (print), 3-642-34179-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34179-3.

Hsu:1991:IAI

[HL91] Wen-Lian Hsu and R. C. T.

Lee, editors. ISA ’91 algorithms: 2nd International Symposium on Algorithms, Taipei, Republic of China, December 16–18, 1991: proceedings, volume 557 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54945-5 (Berlin), 0-387-54945-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .I575 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0557.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54945-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=557.

Hui:2000:IAI

[HL00] Lucas Chi-Kwong Hui and Dik Lun Lee, editors. Internet applications: 5th International Computer Science Conference, ICSC ’99, Hong Kong, China, December 13–15, 1999: Proceedings, volume 1749 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-66903-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1749. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1749.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66903-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1749.

Heery:2004:RAT

[HL04] Rachel Heery and Elizabeth Lyon, editors. Research and advanced technology for digital libraries: 8th European conference, ECDL 2004, Bath, UK, September 12–17, 2004: proceedings, volume 3232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23013-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4080 .E28 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23013-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3232;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100389.

Harder:2005:DMC

[HL05] Theo Härder and Wolfgang Lehner, editors. Data Management in a Connected World: Essays Dedicated to Hartmut Wedekind on the Occasion of His 70th Birthday, volume 3551 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26295-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D356 2005.

URL http://www.springerlink.
com/content/978-3-540-
26295-4;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3551.

Hitzler:2010:WRR

[HL10] Pascal Hitzler and Thomas Lukasiewicz, editors. Web Reasoning and Rule Systems: Fourth International Conference, RR 2010, Bressanone/ Brixen, Italy, September 22–24, 2010. Proceedings, volume 6333 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15917-6 (print), 3-642-15918-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15918-3.

Hao:2012:AEI

[HLC^{+}12] Jin-Kao Hao, Pierrick Legrand, Pierre Collet, Nicolas Monmarché, Evelyne Lutton, and Marc Schoenauer, editors. Artificial Evolution: 10th International Conference, Evolution Artificielle, EA 2011, Angers, France, October 24–26, 2011, Revised Selected Papers, volume 7401 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35532-3 (print), 3-642-35533-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35533-2.

Hu:2011:BII

[HLCZ11] Bin Hu, Jiming Liu, Lin Chen, and Ning Zhong, editors. Brain Informatics: International Conference, BI 2011, Lanzhou, China, September 7–9, 2011. Proceedings, volume 6889 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23604-9 (print), 3-642-23605-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23605-1.

Hullermeier:2012:SUM

[HLFS12] Eyke Hüllermeier, Sebastian Link, Thomas Fober, and Bernhard Seeger, editors. Scalable Uncertainty Management: 6th International Conference, SUM 2012, Marburg, Germany, September 17– 19, 2012. Proceedings, volume 7520 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33361-3 (print), 3-642-33362-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33362-0.

Lee:2009:FGI

[hLhKcFŚ09] Young hoon Lee, Tai hoon Kim, Wai chi Fang, and Dominik Ślezak, editors. Future Generation Information Technology: First International Conference, FGIT 2009, Jeju Island, Korea, December 10–12,2009. Proceedings, volume 5899 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10508-4 (print), 3-642-10509-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10509-8.

Huang:2006:CIB

[HLI06a] De-Shuang Huang, Kang Li, and George William Irwin, editors. Computational Intelligence and Bioinformatics: International Conference on Intelligent Computing, ICIC 2006, Kunming, China, August 16–19, 2006. Proceedings, Part III, volume 4115 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37277-6 (print), 3-540-37282-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37282-0.

Huang:2006:CII

[HLI06b] De-Shuang Huang, Kang Li, and George William Irwin, editors. Computational Intelligence: International Conference on Intelligent Computing, ICIC 2006, Kunming, China, August 16– 19, 2006. Proceedings, Part II, volume 4114 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37274-1 (print), 3-540-37275-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37275-2.

Huang:2006:ICI

[HLI06c] De-Shuang Huang, Kang Li, and George William Irwin, editors. Intelligent Computing: International Conference on Intelligent Computing, ICIC 2006, Kunming, China, August 16– 19, 2006. Proceedings, Part I, volume 4113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37271-7 (print), 3-540-37273-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37273-8.

He:2012:HIS

[HLKX12] Jing He, Xiaohui Liu, Elizabeth A. Krupinski, and Guandong Xu, editors. Health Information Science: First International Conference, HIS 2012, Beijing, China, April 8–10, 2012. Proceedings, volume 7231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29360-3 (print), 3-642-29361-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29361-0.

Havinga:2006:SSC

[HLMW06] Paul Havinga, Maria Lijding, Nirvana Meratnia, and Maarten Wegdam, editors. Smart Sensing and Context: First European Conference, EuroSSC 2006 Enschede, Netherlands, October 25–27, 2006 Proceedings, volume 4272 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47842-6 (print), 3-540-47845-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47845-4.

Hao:1998:AET

[HLR^{+}98] J.-K. (Jin-Kao) Hao, E. (Evelyne) Lutton, E. (Edmund) Ronald, M. (Marc) Schoenauer, and D. (Dominique) Snyers, editors. Artificial evolution: Third European Conference, AE ’97, Nîmes, France, October 22–24, 1997: Selected Papers, volume 1363 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64169-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618.A217 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1363.htm;
 http://www.springerlink.
com/content/978-3-540-
64169-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1363.

Haring:2000:PEO

[HLR00] Günter Haring, Christoph Lindemann, and Martin Reiser, editors. Performance Evaluation: Origins and Directions, volume 1769 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67193-5. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 527 pp. LCCN QA76.9.E94 P45 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1769.htm;
 http://www.springerlink.
com/content/978-3-540-

67193-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1769.

Hameurlain:2011:DESa

[HLSZ11a] Abdelkader Hameurlain, Stephen W. Liddle, Klaus-Dieter Schewe, and Xiaofang Zhou, editors. Database and Expert Systems Applications: 22nd International Conference, DEXA 2011, Toulouse, France, August 29 — September 2, 2011. Proceedings, Part I, volume 6860 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23087-3 (print), 3-642-23088-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23088-2.

Hameurlain:2011:DESb

[HLSZ11b] Abdelkader Hameurlain, Stephen W. Liddle, Klaus-Dieter Schewe, and Xiaofang Zhou, editors. Database and Expert Systems Applications: 22nd International Conference, DEXA 2011, Toulouse, France, August 29 — September 2, 2011, Proceedings, Part II, volume 6861 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23090-3 (print), 3-642-23091-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23091-2.

Heisel:2004:CSR

[HLW04] Maritta Heisel, Peter Liggesmeyer, and Stefan Wittmann, editors. Computer Safety, Reliability, and Security: 23rd International Conference, Safecomp 2004, Potsdam, Germany, September 21– 24, 2004: Proceedings, volume 3219 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23176-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.73 .I58 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3219.
htm;   http://www.springerlink.
com/content/978-3-540-
23176-9;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3219;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100227.

Holldobler:2008:LAI

[HLW08] Steffen Hölldobler, Carsten Lutz, and Heinrich Wansing, editors. Logics in Artificial Intelligence: 11th European Conference, JELIA 2008, Dresden, Germany, September 28–October 1, 2008. Proceedings, volume 5293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87802-5 (print), 3-540-87803-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87803-2.

Habermann:1987:SDA

[HM87] A. Nico Habermann and U. Montanari, editors. System development and Ada, volume 275 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18341-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 W661 1986.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0275.
htm;   http://www.springerlink.
com/content/978-0-387-
18341-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=275.

Contents: Design of the rational environment / James E. Archer, Jr.

— The PCTE initiative / Ferdinando Gallo — Engineering VAX Ada for a multi-language programming environment / Charles Z. Mitchell — The Ada environment — a personal view / Vic Stenning — Knowledgebased software development from requirements to code / Stephen J. Westfold, Lawrence Z. Markosian, William A. Brew — SMoLCS approach to the formal semantics of programming languages / Egidio Astesiano, Gianna Reggio — Project graphs and Meta-programs / Dines Bjorner — Software development based on formal methods / Cliff B. Jones — Integration of program construction and verification / Bernd Krieg-Bruckner

— Ada compiler validation / John B. Goodenough — The Software Engineering Institute at Carnegie Mellon University / A. Nico Habermann — Task sequencing language for specifying distributed Ada systems / David C. Luckham … [et al.].

Hagiya:1994:TAC

[HM94] Masami Hagiya and John C. Mitchell, editors. Theoretical aspects of computer software: International Symposium TACS ’94, Sendai, Japan, April 19–22, 1994: proceedings, volume 789 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-57887-0 (New York), 3-540-57887-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .I568 1994. DM156.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0789.
htm;   http://www.springerlink.
com/content/978-0-387-
57887-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=789.

Hirschberg:1996:CPM

[HM96] Dan Hirschberg and Gene Myers, editors. Combinatorial pattern matching: 7th Annual Symposium, CPM 96, Laguna Beach, California, June 10–12, 1996: proceedings, volume 1075 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61258-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1075.
htm;   http://www.springerlink.
com/content/978-3-540-
61258-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1075.

Husbands:1998:ERF

[HM98] Philip Husbands and Jean-Arcady Meyer, editors. Evolutionary robotics: First European Workshop, EvoRobot’98: Paris, France, April 16–17, 1998: Proceedings, volume 1468 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64957-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ211.37 .E96 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1468.htm;
 http://www.springerlink.
com/content/978-3-540-
64957-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1468.

Honsell:2001:FSS

[HM01] Furio Honsell and Marino Miculan, editors. Foundations of software science and computation structures: 4th International Conference, FOS-SACS 2001, held as part of the Joint European Conference on Theory and Practice of Software, ETAPS 2001, Genova, Italy, April 2–6, 2001: Proceedings, volume 2030 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41864-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2030; QA76.751 .F67 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2030.htm;
 http://www.springerlink.
com/content/978-3-540-
41864-1;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2030.

Hernandez:2002:OOT

[HM02] Juan Hernández and Ana M. D. Moreira, editors. Object-Oriented Technology. ECOOP 2002 Workshop Reader: ECOOP 2002 Workshops and Posters, Málaga, Spain, June 10–14, 2002. Proceedings, volume 2548 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00233-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35 E33 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2548.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00233-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2548.

Henninger:2003:ALS

[HM03] Scott Henninger and Frank Maurer, editors. Advances in Learning Software Organizations: 4th International Workshop, LSO 2002, Chicago, IL, USA, August 6, 2002: Revised Papers, volume 2640 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20591-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 L76 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2640.htm;
 http://www.springerlink.
com/content/978-3-540-
20591-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2640;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94220.

Hu:2004:FMC

[HM04] Alan J. (Alan John) Hu and Andrew K. Martin, editors. Formal methods in computer-aided design: 5th international confrence [i.e., conference], FMCAD 2004, Austin, Texas, USA, November 15–17, 2004: proceedings, volume 3312 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23738-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.65 .F53 2004. URL http://

springerlink.metapress.

                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3312;   http:
                                        //www.springerlink.
                                        com/content/978-3-540-
                                        23738-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3312;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b102264.

Hoos:2005:TAS

[HM05a] Holger H. Hoos and David G. Mitchell, editors. Theory and Applications of Satisfiability Testing: 7th International Conference, SAT 2004, Vancouver, BC, Canada, May 10–13, 2004, Revised Selected Papers, volume 3542 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27829-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.3 .S365 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        27829-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3542.

Hurd:2005:TPH

[HM05b] Joe Hurd and Tom Melham, editors. Theorem proving in higher order logics: 18th international conference, TPHOLs 2005, Oxford, UK, August 22–25, 2005: proceedings, volume 3603 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28372-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T655 2005.

URL http://www.springerlink.
com/content/978-3-540-
28372-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3603.

Halperin:2008:AEA

[HM08] Dan Halperin and Kurt Mehlhorn, editors. Algorithms — ESA 2008: 16th annual European symposium, Karlsruhe, Germany, September 15– 17, 2008, proceedings, volume 5193 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. ISBN 3-540-87744-4, 3-540-87743-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 2008.

URL http://www.springerlink.
com/content/978-3-540-
87743-1/;   http://
www.springerlink.com/

                                        content/978-3-540-87744-
                                        8.

Holzinger:2009:HUI

[HM09] Andreas Holzinger and Klaus Miesenberger, editors. HCI and Usability for e-Inclusion: 5th Symposium of the Workgroup Human-Computer Interaction and Usability Engineering of the Austrian Computer Society, USAB 2009, Linz, Austria, November 9– 10, 2009 Proceedings, volume 5889 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10307-3 (print), 3-642-10308-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10308-7.

Habernal:2011:TSD

[HM11a] Ivan Habernal and Václav Matoušek, editors. Text, Speech and Dialogue: 14th International Conference, TSD 2011, Pilsen, Czech Republic, September 1–5, 2011. Proceedings, volume 6836 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23537-9 (print), 3-642-23538-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23538-2.

Hage:2011:IAF

[HM11b] Jurriaan Hage and Marco T.

Morazán, editors. Implementation and Application of Functional Languages: 22nd International Symposium, IFL 2010, Alphen aan den Rijn, The Netherlands, September 1–3, 2010, Revised Selected Papers, volume 6647 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24275-8 (print), 3-642-24276-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24276-2.

Hsu:2011:MTP

[HM11c] Ching-Hsien Hsu and Victor Malyshkin, editors. Methods and Tools of Parallel Programming Multicomputers: Second Russia-Taiwan Symposium, MTPP 2010, Vladivostok, Russia, May 16–19, 2010, Revised Selected Papers, volume 6083 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-14821-2 (print), 3-642-14822-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14822-4.

Hao:2012:ECC

[HM12a] Jin-Kao Hao and Martin Middendorf, editors. Evolutionary Computation in Combinatorial Optimization: 12th European Conference, EvoCOP 2012, Málaga, Spain, April 11–13, 2012. Proceedings, volume 7245 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29123-6 (print), 3-642-29124-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29124-1.

Hawblitzel:2012:CPP

[HM12b] Chris Hawblitzel and Dale Miller, editors. Certified Programs and Proofs: Second International Conference, CPP 2012, Kyoto, Japan, December 13–15, 2012. Proceedings, volume 7679 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35307-X (print), 3-642-35308-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35308-6.

Hu:2012:CVM

[HM12c] Shi-Min Hu and Ralph R.

Martin, editors. Computational Visual Media: First International Conference, CVM 2012, Beijing, China, November 8–10, 2012. Proceedings, volume 7633 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34262-0 (print), 3-642-34263-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        34263-9.

Hubner:2009:COI

[HMBD09] Jomi Fred Hübner, Eric Matson, Olivier Boissier, and Virginia Dignum, editors. Coordination, Organizations, Institutions and Norms in Agent Systems IV: COIN 2008 International Workshops, COIN@AAMAS 2008, Estoril, Portugal, May 12, 2008. COIN@AAAI 2008, Chicago, USA, July 14, 2008. Revised Selected Papers, volume 5428 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00442-3 (print), 3-642-00443-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00443-8.

Huo:2006:CSL

[HMCL06] Qiang Huo, Bin Ma, Eng-Siong Chng, and Haizhou Li, editors. Chinese Spoken Language Processing: 5th International Symposium, ISC-SLP 2006, Singapore, December 13–16, 2006. Proceedings, volume 4274 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49665-3 (print), 3-540-49666-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49666-3.

Horlait:2003:MAT

[HMG03] Eric Horlait, Thomas Magedanz, and Roch H. Glitho, editors. Mobile Agents for Telecommunication Applications: 5th International Workshop, MATA 2003, Marrakech, Morocco, October 8–10, 2003: Proceedings, volume 2881 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20298-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5101.A1. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2881.htm;
 http://www.springerlink.
com/content/978-3-540-
20298-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2881;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b93926.

Huang:2012:ICTb

[HMJG12] De-Shuang Huang, Jianhua Ma, Kang-Hyun Jo, and M. Michael Gromiha, editors. Intelligent Computing Theories and Applications: 8th International Conference, ICIC 2012, Huangshan, China, July 25–29, 2012. Proceedings, volume 7390 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31575-5 (print), 3-642-31576-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31576-3.

Higuchi:1997:ESB

[HML97] Tetsuya Higuchi, Iwata Masaya, and Weixin Liu, editors. Evolvable systems: from biology to hardware: first international conference, ICES96, Tsukuba, Japan, October 7–8, 1996: proceedings, volume 1259 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63173-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .I57 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1259.htm;
http://www.springerlink.
com/content/978-3-540-
63173-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1259.

Herrlich:2012:ECI

[HMM12] Marc Herrlich, Rainer Malaka, and Maic Masuch, editors. Entertainment Computing — ICEC 2012: 11th International Conference, ICEC 2012, Bremen, Germany, September 26– 29, 2012. Proceedings, volume 7522 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33541-1 (print), 3-642-33542-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33542-6.

Horacek:2010:NLP

[HMMW10] Helmut Horacek, Elisabeth Métais, Rafael Muñoz, and Magdalena Wolska, editors. Natural Language Processing and Information Systems: 14th International Conference on Applications of Natural Language to Information Systems, NLDB 2009, Saarbrücken, Germany, June 24–26, 2009. Revised Papers, volume 5723 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12549-2 (print), 3-642-12550-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12550-8.

Hlavicka:1999:DCE

[HMP99] Jan Hlavička, Erik Maehle, and András Pataricza, editors. Dependable computing—EDCC-3: Third European Dependable Computing Conference, Prague, Czech Republic, September 15–17, 1999: Proceedings, volume 1667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66483-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1667; QA76.9.F38 E95 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1667.
htm;   http://www.springerlink.
com/content/978-3-540-
66483-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1667.

Havelund:2008:MCS

[HMP08] Klaus Havelund, Rupak Majumdar, and Jens Palsberg, editors. Model Checking Software: 15th International SPIN Workshop, Los Angeles, CA, USA, August 10–12, 2008 Proceedings, volume 5156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85113-5 (print), 3-540-85114-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85114-1.

Houben:2009:UMA

[HMPZ09] Geert-Jan Houben, Gord McCalla, Fabio Pianesi, and Massimo Zancanaro, editors. User Modeling, Adaptation, and Personalization: 17th International Conference, UMAP 2009, formerly UM and AH, Trento, Italy, June 22–26, 2009. Proceedings, volume 5535 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02246-4 (print), 3-642-02247-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02247-0.

Hacid:2005:FIS

[HMRT05] Mohand-Said Hacid, Neil V. Murray, Zbigniew W. Raś, and Shusaku Tsumoto, editors. Foundations of intelligent systems: 15th international symposium, ISMIS 2005, Saratoga Springs, NY, USA, May 25–28, 2005, proceedings, volume 3488 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25878-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25878-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3488;
 http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b136302.

Hancock:2009:MSX

[HMS09] Edwin R. Hancock, Ralph R. Martin, and Malcolm A. Sabin, editors. Mathematics of Surfaces XIII: 13th IMA International Conference York, UK, September 7–9, 2009 Proceedings, volume 5654 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03595-7 (print), 3-642-03596-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03596-8.

Hutter:2004:SPC

[HMSU04] Dieter Hutter, Günter Müller, Werner Stephan, and Markus Ullmann, editors. Security in Pervasive Computing: First International Conference, Boppard, Germany, March 12–14, 2003: Revised Papers, volume 2802 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20887-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5915 .I58 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2802.
htm;   http://www.springerlink.
com/content/978-3-540-
20887-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2802;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95124.

Hatcliff:1999:PEP

[HMT99] John Hatcliff, Torben A. Mogensen, and P. Thiemann, editors. Partial evaluation: practice and theory: DIKU 1998 International Summer School, Copenhagen, Denmark, June 29–July 10, 1998, volume 1706 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66710-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E93 D55 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1706.htm;
 http://www.springerlink.

com/content/978-3-540-
66710-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1706.

Hameurlain:2010:DMG

[HMT10a] Abdelkader Hameurlain, Franck Morvan, and A. Min Tjoa, editors. Data Management in Grid and Peer-to-Peer Systems: Third International Conference, Globe 2010, Bilbao, Spain, September 1–2, 2010. Proceedings, volume 6265 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15107-8 (print), 3-642-15108-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15108-8.

Hanrot:2010:ANT

[HMT10b] Guillaume Hanrot, François Morain, and Emmanuel Thomé, editors. Algorithmic Number Theory: 9th International Symposium, ANTS-IX, Nancy, France, July 19– 23, 2010. Proceedings, volume 6197 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14517-5 (print), 3-642-14518-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14518-6.

Hull:2010:BPM

[HMT10c] Richard Hull, Jan Mendling, and Stefan Tai, editors. Business Process Management: 8th International Conference, BPM 2010, Hoboken, NJ, USA, September 13– 16, 2010. Proceedings, volume 6336 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15617-7 (print), 3-642-15618-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15618-2.

Hart:2010:AIS

[HMTH10] Emma Hart, Chris McEwan, Jon Timmis, and Andy Hone, editors. Artificial Immune Systems: 9th International Conference, ICARIS 2010, Edinburgh, UK, July 26– 29, 2010. Proceedings, volume 6209 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14546-9 (print), 3-642-14547-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14547-6.

Hariharan:2001:FTF

[HMV01] Ramesh Hariharan, Madhavan Mukund, and V. Vinay, editors. FST TCS 2001: Foundations of Software Technology and Theoretical Computer Science: 21st Conference, Bangalore, India, December 13–15, 2001. Proceedings, volume 2245 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43002-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F68 2001; QA267.A1 L43 no.2245. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2245.htm;
 http://www.springerlink.
com/content/978-3-540-
43002-5;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2245.

Helal:2008:SHH

[HMW^{+}08] Sumi Helal, Simanta Mitra, Johnny Wong, Carl K. Chang, and Mounir Mokhtari, editors. Smart Homes and Health Telematics: 6th International Conference, ICOST 2008 Ames, IA, USA, June 28–July 2, 2008 Proceedings, volume 5120 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69914-7 (print), 3-540-69916-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69916-3.

Healy:2006:GDI

[HN06] Patrick Healy and Nikola S. Nikolov, editors. Graph drawing: 13th international symposium, GD 2005, Limerick, Ireland, September 12– 14, 2005, revised papers, volume 3843 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31425-3 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31425-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3843.

Hevia:2012:PCL

[HN12] Alejandro Hevia and Gregory Neven, editors. Progress in Cryptology

— LATINCRYPT 2012: 2nd International Conference on Cryptology and Information Security in Latin America, Santiago, Chile, October 7–10, 2012. Proceedings, volume 7533 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33480-6 (print), 3-642-33481-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33481-8.

Hong:2008:ACI

[HNF08] Seok-Hee Hong, Hiroshi Nagamochi, and Takuro Fukunaga, editors. Algorithms and Computation: 19th International Symposium, ISAAC 2008, Gold Coast, Australia, December 15–17, 2008. Proceedings, volume 5369 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92181-8 (print), 3-540-92182-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92182-0.

Haakansson:2009:AMA

[HNH^{+}09] Anne Håkansson, Ngoc Thanh Nguyen, Ronald L. Hartung, Robert J. Howlett, and Lakhmi C. Jain, editors. Agent and Multi-Agent Systems: Technologies and Applications: Third KES International Symposium, KES-AMSTA 2009, Uppsala, Sweden, June 3– 5, 2009. Proceedings, volume 5559 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01664-2 (print), 3-642-01665-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01665-3.

Hegarty:2002:DRI

[HNM02] Mary Hegarty, N. Hari Narayanan, and Bernd E. Meyer, editors. Diagrammatic representation and inference: Second International Conference, Diagrams 2002, Callaway Gardens, GA, USA, April 18–20, 2002: Proceedings, volume 2317 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43561-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA90 .D53 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2317.
htm;   http://www.springerlink.
com/content/978-3-540-
43561-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2317.

Horimoto:2012:ANB

[HNP12]  Katsuhisa  Horimoto,

Masahiko Nakatsui, and Nikolaj Popov, editors. Algebraic and Numeric Biology: 4th International Conference, ANB 2010, Hagenberg, Austria, July 31- August 2, 2010, Revised Selected Papers, volume 6479 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28066-8 (print), 3-642-28067-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28067-2.

Huang:2007:AIH

[HNPP07] Thomas S. Huang, Anton Nijholt, Maja Pantic, and Alex Pentland, editors. Artifical Intelligence for Human Computing: ICMI 2006 and IJCAI 2007 International Workshops, Banff, Canada, November 3, 2006, Hyderabad, India, January 6, 2007, Revised Selected and Invited Papers, volume 4451 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72346-3 (print), 3-540-72348-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72348-6.

Hong:2008:GDI

[HNQ08] Seok-Hee Hong, Takao Nishizeki, and Wu Quan, editors. Graph Drawing: 15th International Symposium, GD 2007, Sydney, Australia, September 24–26, 2007. Revised Papers, volume 4875 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77536-6 (print), 3-540-77537-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77537-9.

Hemmje:2005:IPI

[HNR05] Matthias Hemmje, Claudia Niederee, and Thomas Risse, editors. From integrated publication and information systems to information and knowledge environments: essays dedicated to Erich J. Neuhold on the occasion of his 65th birthday, volume 3379 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24551-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xxiv + 319 pp. LCCN TK7882.I6 F76 2005.

URL http://www.springerlink.

                                        com/content/978-3-540-
                                        24551-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3379;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105978.

Havelund:2006:FAS

[HNRW06] Klaus Havelund, Manuel Núñez, Grigore Roşu, and Burkhart Wolff, editors. Formal Approaches to Software Testing and Runtime Verification: First Combined International Workshops, FATES 2006 and RV 2006, Seattle, WA, USA, August 15–16, 2006, Revised Selected Papers, volume 4262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49699-8 (print), 3-540-49703-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49703-5.

Hou:2012:IRT

[HNS^{+}12] Yuexian Hou, Jian-Yun Nie, Le Sun, Bo Wang, and Peng Zhang, editors. Information Retrieval Technology: 8th Asia Information Retrieval Societies Conference, AIRS 2012, Tianjin, China, December 17–19, 2012. Proceedings, volume 7675 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35340-1 (print), 3-642-35341-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35341-3.

Hromkovic:2004:GTC

[HNW04] Juraj Hromković, Manfred Nagl, and Bernhard Westfechtel, editors. Graph-Theoretic Concepts in Computer Science: 30th International Workshop, WG 2004, Bad Honnef, Germany, June 21–23, 2004: Revised papers, volume 3353 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-24132-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5. URL http:

//www.springerlink.
com/content/978-3-540-
24132-4;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3353;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104584.

Hsiang:1998:ACS

[HO98] Jieh Hsiang and Atsushi Ohori, editors. Advances in computing science, ASIAN ’98: 4th Asian Computing Science Conference, Manila, The Philippines, December 8–10, 1998: Proceedings, volume 1538 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65388-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5.A735 1998.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1538.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65388-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1538.

Hagiya:2003:DCI

[HO03] Masami Hagiya and Azuma Ohuchi, editors. DNA computing: 8th International Workshop on DNA Based Computers, DNA8, Sapporo, Japan, June 10– 13, 2002: Revised Papers, volume 2568 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00531-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.887 .I58 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2568.htm;
 http://www.springerlink.
com/content/978-3-540-
00531-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2568.

Also available via the World Wide Web.

Ho:2012:AIVa

[Ho12a] Yo-Sung Ho, editor. Advances in Image and Video Technology: 5th Pacific Rim Symposium, PSIVT 2011, Gwangju, South Korea, November 20–23, 2011, Proceedings, Part I, volume 7087 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25366-0 (print), 3-642-25367-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25367-6.

Ho:2012:AIVb

[Ho12b] Yo-Sung Ho, editor. Advances in Image and Video Technology: 5th Pacific Rim Symposium, PSIVT 2011, Gwangju, South Korea, November 20–23, 2011, Proceedings, Part II, volume 7088 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25345-8 (print), 3-642-25346-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25346-1.

Haake:2007:GDI

[HOC07] Jörg M. Haake, Sergio F. Ochoa, and Alejandra Cechich, editors. Groupware: Design, Implementation, and Use: 13th International Workshop, CRIWG 2007, Bariloche, Argentina, September 16–20, 2007. Proceedings, volume 4715 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74811-3 (print), 3-540-74812-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74812-0.

Haveraaen:1996:RTD

[HOD96] Magne Haveraaen, Olaf Owe, and Ole-Johan Dahl, editors. Recent trends in data type specification: 11th Workshop on Specification of Abstract Data Types, joint with the 8th COM-PASS Workshop, Oslo, Norway, September 19– 23, 1995: Selected Papers, volume 1130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61629-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A23W67 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1130.
htm;   http://www.springerlink.
com/content/978-3-540-
61629-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1130.

Hoffmann:1982:GTA

[Hof82] Christoph M. (Christoph Martin) Hoffmann.

Group-theoretic algorithms and graph isomorphism, volume 136 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11493-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 311 pp. LCCN QA171 .H675 1982. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0136.htm;
 http://www.springerlink.
com/content/978-0-387-
11493-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=136.

Hofestadt:1997:BGC

[Hof97] Ralf Hofestadt, editor. Bioinformatics: German Conference on Bioinformatics, GCB ’96, Leipzig, Germany, September 30–October 2, 1996: Selected Papers, volume 1278 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63370-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH506.G47 1996.

URL http://link.springer-
ny.com/link/service/

series/0558/tocs/t1278.
htm;   http://www.springerlink.
com/content/978-3-540-
63370-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1278.

Hoffmann:2001:AID

[Hof01] Frank Hoffmann, editor. Advances in intelligent data analysis: 4th International Conference, IDA 2001, Cascais, Portugal, September 13–15, 2001 proceedings, volume 2189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42581-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2189; QA276.A1 I28 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2189.htm;
http://www.springerlink.
com/content/978-3-540-
42581-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2189.

Hoffmann:2003:UPS

[Hof03a] Jörg Hoffmann. Utilizing Problem Structure in Planning: a local search approach, volume 2854 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20259-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 251 pp. LCCN QA63 .H64 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2854.htm;
 http://www.springerlink.
com/content/978-3-540-
20259-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2854;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/

b93903. Revised version of doctoral dissertation, Albert-Ludwigs-Universität Freiburg.

Hofmann:2003:TLC

[Hof03b] Martin Hofmann, editor. Typed lambda calculi and applications: 6th International Conference, TLCA 2003, Valencia, Spain, June 2003: Proceedings, volume 2701 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40332-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.5 .I58 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2701.htm;
http://www.springerlink.
com/content/978-3-540-
40332-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2701.

Hofmann:2011:FSS

[Hof11] Martin Hofmann, editor. Foundations of Software Science and Computational Structures: 14th International Conference, FOSSACS 2011, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2011, Saarbrücken, Germany, March 26–April 3, 2011. Proceedings, volume 6604 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19804-X (print), 3-642-19805-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19805-2.

Holldobler:1989:FEL

[Hol89] Steffen Holldobler. Foundations of equational logic programming, volume 353 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51533-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 250 pp. LCCN QA76.63 .H65 1989. US$24.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0353.htm;
 http://www.springerlink.
com/content/978-0-387-
51533-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=353.

Holzinger:2007:HUM

[Hol07] Andreas Holzinger, editor. HCI and Usability for Medicine and Health Care: Third Symposium of the Workgroup Human-Computer Interaction and Usability Engineering of the Austrian Computer Society, USAB 2007 Graz, Austria, November, 22, 2007. Proceedings, volume 4799 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76804-1 (print), 3-540-76805-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76805-0.

Holzinger:2008:HUE

[Hol08] Andreas Holzinger, editor. HCI and Usability for Education and Work: 4th Symposium of the Workgroup Human-Computer Interaction and Usability Engineering of the Austrian Computer Society, USAB 2008, Graz, Austria, November 20–21, 2008. Proceedings, volume 5298 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89349-0 (print), 3-540-89350-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89350-9.

Honary:2001:CCI

[Hon01] Bahram Honary, editor. Cryptography and coding: 8th IMA International Conference, Cirencester, UK, December 17–19, 2001: Proceedings, volume 2260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43026-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .C76 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2260.htm;
 http://www.springerlink.
com/content/978-3-540-
43026-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2260.

Hooman:1991:SCV

[Hoo91] J. (Jozef) Hooman. Specification and compositional verification of realtime systems, volume 558 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54947-1 (New York), 3-540-54947-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 235 pp. LCCN QA76.54 .H66 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0558.
htm;   http://www.springerlink.
com/content/978-0-387-
54947-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=558.

Rev. version of the author’s thesis (Ph. D.)– Eindhoven University of Technology.

Han:1997:ICS

[HOQ97] Yongfei Han, Tatsuaki Okamoto, and Sihan Qing, editors. Information and communications security: first international conference, ICICS ’97, Beijing, China, November 11–13, 1997: proceedings, volume 1334 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63696-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25I554 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1334.htm;
http://www.springerlink.
com/content/978-3-540-
63696-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1334.

Horlait:2000:MAT

[Hor00] Eric Horlait, editor. Mobile Agents for Telecommunication Applications: Second International Workshop, MATA 2000, Paris, France, September 18– 20, 2000. Proceedings, volume 1931 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41069-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1931. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1931.htm;
 http://www.springerlink.
com/content/978-3-540-
41069-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1931.

Horspool:2002:CCI

[Hor02] R. Nigel Horspool, editor. Compiler construction: 11th International Conference, CC 2002 held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2002, Grenoble, France, April 8–12, 2002: Proceedings, volume 2304 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43369-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 2002. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t2304.htm;
http://www.springerlink.
com/content/978-3-540-
43369-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2304.

Horvath:2006:CEF

[Hor06] Zoltán Horváth, editor. Central European Functional Programming School: First Summer School, CEFP 2005, Budapest, Hungary, July 4–15, 2005, Revised Selected Lectures, volume 4164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46843-9 (print), 3-540-46845-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46845-5.

Houbak:1990:SSL

[Hou90] Niels Houbak. SIL — a simulation language: user’s guide, volume 426 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52497-5 (Berlin), 0-387-52497-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). v + 192 pp. LCCN QA76.9.C65 H68 1990. DM33.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0426.htm;
 http://www.springerlink.
com/content/978-3-540-
52497-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=426.

Huguet:1989:AAA

[HP89] L. Huguet and Alain Poli, editors. Applied algebra, algebraic algorithms, and errorcorrecting codes: 5th International Conference, AAECC-5, Menorca, Spain, June 15–19, 1987: proceedings, volume 356 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 3-540-51082-6 (Berlin), 0-387-51082-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 1987. US$32.90. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0356.htm;
 http://www.springerlink.
com/content/978-3-540-
51082-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=356.

Hermenegildo:1994:PLI

[HP94] Manuel Hermenegildo and Jaan Penjam, editors. Programming language implementation and logic programming: 6th International Symposium, PLILP ’94, Madrid, Spain, September 14–16, 1994: proceedings, volume 844 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58402-1 (Berlin), 0-387-58402-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .I5773 1994. DM86.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0844.
htm;   http://www.springerlink.
com/content/978-3-540-
58402-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=844.

Hartel:1995:FPL

[HP95] Pieter H. Hartel and Rinus Plasmeijer, editors. Functional programming languages in education: first International Symposium, FPLE ’95, Nijmegen, the Netherlands, December 4–6, 1995: proceedings, volume 1022 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60675-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .F68 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1022.htm;
 http://www.springerlink.
com/content/978-3-540-
60675-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1022.

Hunter:1998:AUF

[HP98] Anthony Hunter and Simon Parsons. Applications of uncertainty formalisms, volume 1455 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65312-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 474 pp. LCCN Q375 .A66 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1455.
htm;   http://www.springerlink.
com/content/978-3-540-
65312-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1455.

Halbwachs:1999:CAV

[HP99a] Nicolas Halbwachs and Doron Peled, editors. Computer aided verification: 11th International Conference, CAV’99, Trento, Italy, July 6– 10, 1999: Proceedings, volume 1633 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66202-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 C38 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1633.
htm;   http://www.springerlink.
com/content/978-3-540-
66202-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1633.

Hancock:1999:EMM

[HP99b] Edwin R. Hancock and Marcello Pelillo, editors. Energy minimization methods in computer vision and pattern recognition: Second International Workshop, EMM-CVPR ’99, York, UK, July 26–29, 1999: Proceedings, volume 1654 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66294-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1654. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1654.htm;
 http://www.springerlink.
com/content/978-3-540-
66294-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1654.

Hunter:1999:SQA

[HP99c] Anthony Hunter and Simon Parsons, editors. Symbolic and quantitative approaches to reasoning and uncertainty: European Conference, ECSQARU’99, London, UK, July 5–9, 1999: Proceedings, volume 1638 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66131-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q375 .E88 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1638.htm;
 http://www.springerlink.
com/content/978-3-540-
66131-3;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1638.

Haake:2002:GDI

[HP02a] Jorg M. Haake and Jose A. Pino, editors. Groupware — Design, Implementation, and Use: 8th International Workshop, CRIWG 2002, La Serena, Chile, September 1–4, 2002: Proceedings, volume 2440 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44112-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HD66.2 .I58 2002; LB1028.43 I684 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2440.
htm;   http://www.springerlink.
com/content/978-3-540-
44112-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2440.

Hermenegildo:2002:SAI

[HP02b] Manuel Hermenegildo and German Puebla, editors. Static Analysis: 9th International Symposium, SAS 2002, Madrid, Spain, September 17–20, 2002. Proceedings, volume 2477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44235-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2477.htm;
 http://www.springerlink.
com/content/978-3-540-
44235-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2477.

Hermanns:2006:TAC

[HP06] Holger Hermanns and Jens Palsberg, editors. Tools and Algorithms for the Construction and Analysis of Systems: 12th International Conference, TACAS 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 25–April 2, 2006. Proceedings, volume 3920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33056-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
33056-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3920.

Heuser:2009:ACM

[HP09] Carlos Alberto Heuser and Günther Pernul, editors. Advances in Conceptual Modeling — Challenging Perspectives: ER 2009 Workshops Co-MoL, ETheCoM, FP-UML, MOST-ONISW, QoIS, RIGiM, SeCo-GIS, Gramado, Brazil, November 9–12, 2009. Proceedings, volume 5833 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04946-X (print), 3-642-04947-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04947-7.

Haddad:2012:ATP

[HP12] Serge Haddad and Lucia Pomello, editors. Application and Theory of Petri Nets: 33rd International Conference, PETRI NETS 2012, Hamburg, Germany, June 25–29, 2012. Proceedings, volume 7347 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31130-X (print), 3-642-31131-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31131-4.

Houstis:1988:SIC

[HPP88] E. N. Houstis, T. S. Papatheodorou, and C. D. Polychronopoulos, editors. Supercomputing: 1st International Conference, Athens, Greece, June 8–12, 1987: proceedings, volume 297 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-18991-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.297. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0297.htm;
 http://www.springerlink.
com/content/978-0-387-
18991-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=297.

“The conference was organized and sponsored by the Computer Technology Institute (C.T.I.) of Greece”–Pref.

Hess:2006:ANT

[HPP06] Florian Hess, Sebastian Pauli, and Michael Pohst, editors. Algorithmic Number Theory: 7th International Symposium, ANTS-VII, Berlin, Germany, July 23–28, 2006. Proceedings, volume 4076 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36075-1 (print), 3-540-36076-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36076-6.

Herrero:2005:SAG

[HPR05] Pilar Herrero, María S.

Pérez, and Víctor Robles, editors. Scientific applications of grid computing: first international workshop, SAG 2004, Beijing, China, September 20–24, 2004. Revised selected and invited papers, volume 3458 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25810-8 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25810-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3458;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136149.

Hoogeboom:2006:MCI

[HPRS06] Hendrik Jan Hoogeboom, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa, editors. Membrane Computing: 7th International Workshop, WMC 2006, Leiden, The Netherlands, July 17–21, 2006, Revised, Selected, and Invited Papers, volume 4361 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69088-3 (print), 3-540-69090-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69090-0.

Hindriks:2009:PMA

[HPS09] Koen V. Hindriks, Alexander Pokahr, and Sebastian Sardina, editors. Programming Multi-Agent Systems: 6th International Workshop, ProMAS 2008, Estoril, Portugal, May 13, 2008. Revised Invited and Selected Papers, volume 5442 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03277-X (print), 3-642-03278-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03278-3.

Horvath:2008:CEF

[HPSZ08] Zoltán Horváth, Rinus Plasmeijer, Anna Soós, and Viktória Zsók, editors. Central European Functional Programming School: Second Summer School, CEFP 2007, Cluj-Napoca, Romania, June 23–30, 2007, Revised Selected Lectures, volume 5161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88058-5 (print), 3-540-88059-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88059-2.

Hamalainen:2005:ECS

[HPTV05] Timo D. Hämäläinen, Andy D. Pimentel, Jarmo Takala, and Stamatis Vassiliadis, editors. Embedded Computer Systems: Architectures, Modeling, and Simulation: 5th International Workshop, SAMOS 2005, Samos, Greece, July 18–20, 2005. Proceedings, volume 3553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26969-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A73 S27 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        26969-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3553.

Havelund:2000:SMC

[HPV00] Klaus Havelund, John Penix, and Willem Visser, editors. SPIN model checking and software verification: 7th International SPIN Workshop, Stanford, CA, USA, August 30–September 1, 2000: Proceedings, volume 1885 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41030-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 I58 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1885.
htm;   http://www.springerlink.
com/content/978-3-540-
41030-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1885.

Horvath:2010:CEF

[HPZ10] Zoltán Horváth, Rinus Plasmeijer, and Viktória Zsók, editors. Central European Functional Programming School: Third Summer School, CEFP 2009, Budapest, Hungary, May 21–23, 2009 and Komárno, Slovakia, May 25–30, 2009, Revised Selected Lectures, volume 6299 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17684-4 (print), 3-642-17685-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        17685-2.

Herzog:1991:TUL

[HR91] O. Herzog and Claus-Rainer Rollinger. Text understanding in LILOG: integrating computational linguistics and artificial intelligence: final report on the IBM Germany LILOG-project, volume 546 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-54594-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 738 pp. LCCN P98 T49 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0546.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        54594-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=546.

Helary:1995:DAI

[HR95] Jean-Michel Helary and M. Raynal, editors. Distributed algorithms: 9th International Workshop, WDAG ’95, Le-Mont-Saint-Michel, France, September 13–15, 1995, proceedings, volume 972 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60274-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I59 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0972.
htm;   http://www.springerlink.
com/content/978-3-540-
60274-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=972.

Hanus:1996:ALP

[HR96] Michael Hanus and M. (Mario)

Rodriguez Artalejo, editors. Algebraic and logic programming: 5th international conference, ALP ’96, Aachen, Germany, September 25– 27, 1996: proceedings, volume 1139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61735-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .A45844 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1139.
htm;   http://www.springerlink.
com/content/978-3-540-
61735-8;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1139.

Proceedings of the Fifth International Conference on Algebraic and Logic Programming (ALP ’96).

Hellwagner:1999:SSC

[HR99a]  Hermann  Hellwagner

and Alexander Reinefeld. SCI-Scalable coherent interface: architecture and software for highperformance compute clusters, volume 1734 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66696-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xxi + 490 pp. LCCN QA76.88 .S313 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1734.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66696-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1734.

Huntbach:1999:AOP

[HR99b] Matthew M. Huntbach and Graem A. Ringwood. Agent-oriented programming: From Prolog to Guarded Definite Clauses, volume 1630 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66683-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 386 pp. LCCN QA76.64 .H86 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1630.
htm;   http://www.springerlink.
com/content/978-3-540-
66683-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1630.

Halevi:2006:TCT

[HR06] Shai Halevi and Tal Rabin, editors. Theory of Cryptography: Third Theory of Cryptography Conference, TCC 2006, New York, NY, USA, March 4–7, 2006. Proceedings, volume 3876 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32731-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32731-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3876.

Hirschfeld:2008:SSS

[HR08] Robert Hirschfeld and Kim Rose, editors. Self-Sustaining Systems: First Workshop, S3 2008 Potsdam, Germany, May 15– 16, 2008 Revised Selected Papers, volume 5146 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89274-5 (print), 3-540-89275-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89275-5.

Hu:2002:FLP

[HRA02] Zhenjiang Hu and M. Rodríguez-Artalejo, editors. Functional and logic programming: 6th International Symposium, FLOPS 2002, Aizu, Japan, September 15– 17, 2002: Proceedings, volume 2441 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44233-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .F84 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2441.htm;
http://www.springerlink.
com/content/978-3-540-
44233-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2441.

Herkersdorf:2012:ACS

[HRB12] Andreas Herkersdorf, Kay Römer, and Uwe Brinkschulte, editors. Architecture of Computing Systems — ARCS 2012: 25th International Conference, Munich, Germany, February 28 — March 2, 2012. Proceedings, volume 7179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28292-X (print), 3-642-28293-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28293-5.

Harper:2006:ECI

[HRC06] Richard Harper, Matthias Rauterberg, and Marco Combetto, editors. Entertainment Computing

— ICEC 2006: 5th International Conference, Cambridge, UK, September 20–22, 2006. Proceedings, volume 4161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45259-1 (print), 3-540-45261-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45261-4.

Hanbury:2011:MIR

[HRdV11] Allan Hanbury, Andreas Rauber, and Arjen P. de Vries, editors. Multidisciplinary Information Retrieval: Second Information Retrieval Facility Conference, IRFC 2011, Vienna, Austria, June 6, 2011. Proceedings, volume 6653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21352-9 (print), 3-642-21353-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21353-3.

Hainaut:2007:ACM

[HRK^{+}07] Jean-Luc Hainaut, Elke A. Rundensteiner, Markus Kirchberg, Michela Bertolotto, Mathias Brochhausen, Yi-Ping Phoebe Chen, Samira Si-Saïd Cherfi, Martin Doerr, Hyoil Han, Sven Hartmann, Jeffrey Parsons, Geert Poels, Colette Rolland, Juan Trujillo, Eric Yu, and Esteban Zimányie, editors. Advances in Conceptual Modeling — Foundations and Applications: ER 2007 Workshops CMLSA, FP-UML, ONISW, QoIS, RIGiM,SeCoGIS, Auckland, New Zealand, November 5–9, 2007. Proceedings, volume 4802 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76291-4 (print), 3-540-76292-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76292-8.

Hopfe:2010:NLP

[HRM^{+}10] Christina J. Hopfe, Yacine Rezgui, Elisabeth Métais, Alun Preece, and Haijiang Li, editors. Natural Language Processing and Information Systems: 15th International Conference on Applications of Natural Language to Information Systems, NLDB 2010, Cardiff, UK, June 23– 25, 2010. Proceedings, volume 6177 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13880-2 (print), 3-642-13881-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13881-2.

Hanson:1993:MLT

[HRR93] Stephen Jose Hanson, Werner Remmele, and Ronald L. Rivest, editors. Machine learning: from theory to applications: cooperative research at Siemens and MIT, volume 661 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56483-7 (New York), 3-540-56483-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .M322 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0661.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-

56483-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=661.

This volume includes some of the key research papers in the area of machine learning produced at MIT and Siemens.

Hinchey:2006:ICA

[HRR^{+}06] Michael G. Hinchey, Patricia Rago, James L. Rash, Christopher A. Rouff, Roy Sterritt, and Walt Truszkowski, editors. Innovative Concepts for Autonomic and Agent-Based Systems: Second International Workshop on Radical Agent Concepts, WRAC 2005, Greenbelt, MD, USA, September 20–22, 2005. Revised Papers, volume 3825 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69265-7 (print), 3-540-69266-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69266-9.

Horimoto:2008:ABT

[HRRY08]  Katsuhisa  Horimoto,

Georg Regensburger, Markus Rosenkranz, and Hiroshi Yoshida, editors. Algebraic Biology: Third International Conference, AB 2008, Castle of Hagenberg, Austria, July 31–August 2, 2008 Proceedings, volume 5147 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85100-3 (print), 3-540-85101-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85101-1.

Hirsch:2008:CST

[HRSS08] Edward A. Hirsch, Alexander A. Razborov, Alexei Semenov, and Anatol Slissenko, editors. Computer Science — Theory and Applications: Third International Computer Science Symposium in Russia, CSR 2008 Moscow, Russia, June 7–12, 2008 Proceedings, volume 5010 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79708-4 (print), 3-540-79709-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79709-8.

Hinchey:2003:FAA

[HRT^{+}03] Michael G. (Gerard)

Hinchey, James L. Rash, Walter F. Truszkowski, Christopher Rouff, and Diana Gordon-Spears, editors. Formal Approaches to Agent-Based Systems: Second International Workshop, FAABS 2002 Greenbelt, MD, USA, October 29– 31, 2002 Revised Papers, volume 2699 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40665-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ217.5 .F33 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2699.htm;
http://www.springerlink.
com/content/978-3-540-
40665-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2699;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11729.

Hinchey:2005:FAA

[HRTR05] Michael G. (Michael Gerard) Hinchey, James L.

Rash, Walter F. Truszkowski, and Christopher A. Rouff, editors. Formal approaches to agentbased systems: Third International Workshop, FAABS 2004, Greenbelt, MD, USA, April 26–27, 2004: revised selected papers, volume 3228 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24422-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ217.5 .F33 2004. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3228;
 http://www.springerlink.
com/content/978-3-540-
24422-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3228.

Hacid:2002:FIS

[HRZK02] Mohand-Saïd Hacid, Zbigniew W. Raś, Djamel A. Zighed, and Yves Kodratoff, editors. Foundations of Intelligent Systems: 13th International Symposium, IS-MIS 2002, Lyon, France, June 27–29, 2002. Proceedings, volume 2366 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43785-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2366.
htm;   http://www.springerlink.
com/content/978-3-540-
43785-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2366.

Handler:1974:RBE

[HS74] Wolfgang Handler and Peter P. Spies, editors. Rechnerstrukturen und Betriebsprogrammierung, Erlangen, 1970. (German) [Computer Structures and Operating Programming, Erlangen, 1970], volume 13 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 0-387-06815-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6.F3 1970. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0013.
htm;   http://www.springerlink.

                                        com/content/978-0-387-
                                        06815-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=13.

Conference sponsored jointly by the Gesellschaft für Informatik and the Nachrichtentechnische Gesellschaft. English or German.

Hasselmeier:1976:DBS

[HS76] H. Hasselmeier and W. G. Spruth, editors. Data base systems: proceedings, 5th Informatik Symposium, IBM Germany, Bad Homburg v.d.H., September 24– 26, 1975, volume 39 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1976. CO-DEN LNCSD9. ISBN 0-387-07612-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3I521 1975.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0039.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        07612-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=39.

English or German. Sponsored by IBM Germany and the IBM World Trade Corporation.

Hartenstein:1994:FPL

[HS94] Reiner Hartenstein and Michal Z. Servit, editors. Field-programmable logic: architectures, synthesis, and applications: 4th International Workshop on Field-Programmable Logic and Applications, FPL ’94, Prague, Czech Republic, September 7–9, 1994: proceedings, volume 849 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58419-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7869.L6 I56 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0849.
htm;   http://www.springerlink.
com/content/978-0-387-
58419-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=849.

Hermenegildo:1995:PLI

[HS95a] Manuel Hermenegildo and Doaitse Swierstra, editors. Programming languages: implementations, logics and programs: 7th International Symposium, PLILP ’95, Utrecht, The Netherlands, September 20– 22, 1995: proceedings, volume 982 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60359-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .I5773 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0982.htm;
 http://www.springerlink.
com/content/978-3-540-
60359-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=982.

Hertzberger:1995:HPC

[HS95b] Bob Hertzberger and Giuseppe Serazzi, editors. High-performance computing and networking: International Conference and Exhibition, Milan, Italy, May 3– 5, 1995: proceedings, volume 919 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0919.htm;
 http://www.springerlink.

com/content/978-3-540-
59393-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=919.

Hlavac:1995:CAI

[HS95c] Vaclav Hlavac and Radim Sara, editors. Computer analysis of images and patterns: 6th International Conference, CAIP ’95, Prague Czech Republic, September 6–8, 1995: proceedings, volume 970 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60268-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .I53 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0970.htm;
http://www.springerlink.
com/content/978-3-540-
60268-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=970.

Hertzberger:1997:HPC

[HS97] Bob Hertzberger and Peter Sloot, editors. Highperformance computing and networking: international conference and exhibition, Vienna, Austria, April 28–30, 1997: proceedings, volume 1225 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62898-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H526 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1225.htm;
 http://www.springerlink.
com/content/978-3-540-
62898-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1225.

Henzinger:1998:HSC

[HS98a] T. A. (Thomas A.) Henzinger and Shankar (Sosale Shankara) Sastry, editors. Hybrid systems: computation and control: First International Workshop, HSCC’98, Berkeley, California, USA, April 1998: Proceedings, volume 1386 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64358-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38.H73 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1386.
htm;   http://www.springerlink.

com/content/978-3-540-
64358-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1386.

Honavar:1998:GII

[HS98b] Vasant Honavar and Giora Slutzki, editors. Grammatical inference: 4th international colloquium, ICGI-98, Ames, Iowa, USA, July 12– 14, 1998: Proceedings, volume 1433 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64776-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .G66 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1433.htm;
http://www.springerlink.
com/content/978-3-540-
64776-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1433.

Hromkovic:1998:GTC

[HS98c] Juraj Hromkovic and Ondrej Sykora, editors. Graph theoretic concepts in computer science: 24th International Workshop, WG’98, Smolenice Castle, Slovak Republic, June 18–20: Proceedings, volume 1517 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65195-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G7172 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1517.htm;
 http://www.springerlink.
com/content/978-3-540-
65195-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1517.

Hutchison:1998:MAS

[HS98d] David Hutchison and Ralf Schafer, editors. Multimedia applications, services, and techniques: ECMAST ’98, Third European Conference, Berlin, Germany, May 26–28, 1998: Proceedings, volume 1425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64594-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1425.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1425.
htm;   http://www.springerlink.

com/content/978-3-540-
64594-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1425.

Huijsmans:1999:VII

[HS99] Dionysius P. Huijsmans and Arnoldus Wilhelmus Maria Smeulders, editors. Visual information and information systems: Third International Conference, VISUAL ’99, Amsterdam, The Netherlands, June 2–4, 1999: Proceedings, volume 1614 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66079-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.I6 I675 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1614.htm;
http://www.springerlink.
com/content/978-3-540-
66079-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1614.

He:2000:ACS

[HS00] Jifeng He and Masahiko Sato, editors. Advances in computing science — ASIAN 2000: 6th Asian Computing Science Conference, Penang, Malaysia, November 2000: Proceedings, volume 1961 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41428-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .A735 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1961.htm;
 http://www.springerlink.
com/content/978-3-540-
41428-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1961.

Hermanns:2002:PAP

[HS02] Holger Hermanns and Roberto Segala, editors. Process Algebra and Probabilistic Methods. Performance Modeling and Verification: Second Joint International Workshop PAPM-PROBMIV 2002, Copenhagen, Denmark, July 25–26, 2002. Proceedings, volume 2399 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43913-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .P36 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2399.htm;
http://www.springerlink.
com/content/978-3-540-
43913-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2399.

Hunt:2003:CAV

[HS03] Warren A. Hunt and Fabio Somenzi, editors. Computer aided verification: 15th International Conference, CAV 2003, Boulder, CO, USA, July 8–12, 2003: Proceedings, volume 2725 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40524-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 C38 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2725.htm;
http://www.springerlink.
com/content/978-3-540-
40524-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2725.

Hankin:2005:SAI

[HS05a] Chris Hankin and Igor Siveroni, editors. Static analysis: 12th international symposium, SAS 2005, London, UK, September 7–9, 2005, proceedings, volume 3672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28584-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28584-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3672.

Hutter:2005:MMR

[HS05b] Dieter Hutter and Werner Stephan. Mechanizing mathematical reasoning: essays in honor of Joerg H. Siekmann on the occasion of his 60th birthday, volume 2605 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25051-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25051-7;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2605.

Harders:2006:BST

[HS06] Matthias Harders and Gábor Székely, editors. Biomedical Simulation: Third International Symposium, ISBMS 2006, Zurich, Switzerland, July 10–11, 2006. Proceedings, volume 4072 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36009-3 (print), 3-540-36010-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36010-0.

Hammerli:2007:DIM

[HS07] Bernhard M. Hämmerli and Robin Sommer, editors. Detection of Intrusions and Malware, and Vulnerability Assessment: 4th International Conference, DIMVA 2007 Lucerne, Switzerland, July 12– 13, 2007 Proceedings, volume 4579 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73613-1 (print), 3-540-73614-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73614-1.

Harrison:2008:CSR

[HS08a] Michael D. Harrison and Mark-Alexander Sujan, editors. Computer Safety, Reliability, and Security: 27th International Conference, SAFECOMP 2008 Newcastle upon Tyne, UK, September 22–25, 2008 Proceedings, volume 5219 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87697-9 (print), 3-540-87698-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87698-4.

Hausheer:2008:RNS

[HS08b] David Hausheer and Jürgen Schönwälder, editors. Resilient Networks and Services: Second International Conference on Autonomous Infrastructure, Management and Security, AIMS 2008 Bremen, Germany, July 1–3, 2008 Proceedings, volume 5127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70586-4 (print), 3-540-70587-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70587-1.

Hummel:2008:SOS

[HS08c] Karin Anna Hummel and James P. G. Sterbenz, editors. Self-Organizing Systems: Third International Workshop, IW-SOS 2008, Vienna, Austria, December 10–12, 2008. Proceedings, volume 5343 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92156-7 (print), 3-540-92157-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92157-8.

Holzinger:2011:IQH

[HS11] Andreas Holzinger and Klaus-Martin Simonic, editors. Information Quality in e-Health: 7th Conference of the Workgroup Human-Computer Interaction and Usability Engineering of the Austrian Computer Society, USAB 2011, Graz, Austria, November 25– 26, 2011. Proceedings, volume 7058 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25363-6 (print), 3-642-25364-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25364-5.

Hadjiantonis:2012:TES

[HS12a] Antonis M. Hadjiantonis and Burkhard Stiller, editors. Telecommunication Economics: Selected Results of the COST Action ISO605 Econ@Tel, volume 7216 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30381-1 (print), 3-642-30382-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30382-1.

Hamadi:2012:LIO

[HS12b] Youssef Hamadi and Marc Schoenauer, editors. Learning and intelligent optimization: 6th International Conference, LION 6, Paris, France, January 16–20, 2012, Revised selected papers, volume 7219 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34413-5 (e-book). LCCN Q325.5 .L56 2012. URL http:

//www.springerlink.
com/content/978-3-642-
34413-8.

Huang:1996:LCP

[HSB^{+}96] C.-H. (Chua-Huang) Huang, P. Sadayappan, Utpal Banerjee, David Gelernter, Alexandru Nicolau, and David A. Padua, editors. Languages and compilers for parallel computing: 8th International Workshop, LCPC 95, Columbus, Ohio, USA, August 10–12, 1995: proceedings, volume 1033 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60765-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1033.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60765-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1033.

Holscher:2010:SCV

[HSB^{+}10] Christoph Hölscher, Thomas F. Shipley, Marta Olivetti Belardinelli, John A. Bateman, and Nora S. Newcombe, editors. Spatial Cognition VII: International Conference, Spatial Cognition 2010, Mt. Hood/Portland, OR, USA, August 15–19, 2010. Proceedings, volume 6222 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14748-8 (print), 3-642-14749-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14749-4.

Hornby:2008:ESB

[HSH08] Gregory S. Hornby, Lukáš Sekanina, and Pauline C. Haddow, editors. Evolvable Systems: From Biology to Hardware: 8th International Conference, ICES 2008, Prague, Czech Republic, September 21–24, 2008. Proceedings, volume 5216 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85856-3 (print), 3-540-85857-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85857-7.

Hsiang:1995:RTA

[Hsi95] Jieh Hsiang, editor.

Rewriting techniques and applications: 6th international conference, RTA-95, Kaiserslautern, Germany, April 5–7, 1995: proceedings, volume 914 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59200-8 (Berlin), 0-387-59200-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I555 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0914.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        59200-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=914.

Hamprecht:2007:PRD

[HSJ07] Fred A. Hamprecht, Christoph Schnörr, and Bernd Jähne, editors. Pattern Recognition: 29th DAGM Symposium, Heidelberg, Germany, September 12–14, 2007. Proceedings, volume 4713 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74933-0 (print), 3-540-74936-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74936-3.

Ho:2009:DWI

[HSKB09] Anthony T. S. Ho, Yun Q. Shi, H. J. Kim, and Mauro Barni, editors. Digital Watermarking: 8th International Workshop, IWDW 2009, Guildford, UK, August 24–26, 2009. Proceedings, volume 5703 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03687-2 (print), 3-642-03688-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03688-0.

Horn:1999:AIM

[HSL^{+}99] Werner Horn, Yuval Shahar, Greger Lindberg, Steen Andreassen, and Jeremy Wyatt, editors. Artificial intelligence in medicine: Joint European Conference on Artificial Intelligence in Medicine and Medical Decision Making, AIMDM’99, Aalborg, Denmark, June 20–24, 1999: Proceedings, volume 1620 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66162-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R859.7.A78 J65 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1620.
htm;   http://www.springerlink.
com/content/978-3-540-
66162-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1620.

Huang:2006:CVH

[HSL06] Thomas S. Huang, Nicu Sebe, and Michael S. Lew, editors. Computer Vision in Human-Computer Interaction: ECCV 2006 Workshop on HCI, Graz, Austria, May 13, 2006. Proceedings, volume 3979 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34202-8 (softcover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34202-1.

Huet:2009:AMM

[HSMPA09] Benoit Huet, Alan Smeaton, Ketan Mayer-Patel, and Yannis Avrithis, editors. Advances in Multimedia Modeling: 15th International Multimedia Modeling Conference, MMM 2009, Sophia-Antipolis, France, January 7–9, 2009. Proceedings., volume 5371 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-92891-X (print), 3-540-92892-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
92892-8.

Heyden:2002:CVEb

[HSNJ02a] A. Heyden, G. Sparr, M. Nielsen, and P. Johansen, editors. Computer Vision—ECCV 2002: 7th European Conference on Computer Vision, Copenhagen, Denmark, May 28–31, 2002. Proceedings, Part II, volume 2351 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43744-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E97 2002 pt.1-4 (2002). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2351.htm;
http://www.springerlink.
com/content/978-3-540-
43744-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2351.

Heyden:2002:CVEc

[HSNJ02b] A. Heyden, G. Sparr, M. Nielsen, and P. Johansen, editors. Computer Vision—ECCV 2002: 7th European Conference on Computer Vision, Copenhagen, Denmark, May 28–31, 2002. Proceedings, Part III, volume 2352 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43746-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E97 2002 pt.1-4 (2002). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2352.htm;
 http://www.springerlink.
com/content/978-3-540-
43746-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2352.

Heyden:2002:CVEd

[HSNJ02c] A. Heyden, G. Sparr, M. Nielsen, and P. Johansen, editors. Computer Vision—ECCV 2002: 7th European Conference on Computer Vision, Copenhagen, Denmark, May 28–31, 2002. Proceedings, Part IV, volume 2353 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43748-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E97 2002 pt.1-4 (2002). URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t2353.htm;
http://www.springerlink.
com/content/978-3-540-
43748-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2353.

Heyden:2002:CVEa

[HSNJ02d] Anders Heyden, Gunnar Sparr, Mads Nielsen, and Peter Johansen, editors. Computer Vision—ECCV 2002: 7th European Conference on Computer Vision, Copenhagen, Denmark, May 28–31, 2002. Proceedings, Part I, volume 2350 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43745-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E97 2002 pt.1-4 (2002). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2350.htm;
http://www.springerlink.
com/content/978-3-540-
43745-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2350.

Herrmann:2012:AAR

[HSP^{+}12] Guido Herrmann, Matthew Studley, Martin Pearson, Andrew Conn, Chris Melhuish, Mark Witkowski, Jong-Hwan Kim, and Prahlad Vadakkepat, editors. Advances in Autonomous Robotics: Joint Proceedings of the 13th Annual TAROS Conference and the 15th Annual FIRA RoboWorld Congress, Bristol, UK, August 20–23, 2012, volume 7429 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32526-2 (print), 3-642-32527-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32527-4.

Hlawiczka:1996:DCE

[HSS96] Andrzej Hlawiczka, Joao Gabriel Silva, and L. (Luca) Simoncini, editors. Dependable computing — EDCC-2: second European Dependable Computing Conference, Taormina, Italy, October 2–4, 1996: proceedings, volume 1150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61772-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F38 E95 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1150.
htm;   http://www.springerlink.
com/content/978-3-540-
61772-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1150.

Hightower:2007:LCA

[HSS07] Jeffrey Hightower, Bernt Schiele, and Thomas Strang, editors. Locationand Context-Awareness: Third International Symposium, LoCA 2007, Oberpfaffenhofen, Germany, September 20– 21, 2007. Proceedings, volume 4718 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75159-9 (print), 3-540-75160-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75160-1.

Hu:2012:CLT

[HSSV12] Hao Hu, Xiaoning Shi, Robert Stahlbock, and Stefan Voß, editors. Computational Logistics: Third International Conference, ICCL 2012, Shanghai, China, September 24–26, 2012. Proceedings, volume 7555 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33586-1 (print), 3-642-33587-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33587-7.

Helleseth:2005:STA

[HSSY05] Tor Helleseth, Dilip Sarwate, Hong-Yeop Song, and Kyeongcheol Yang, editors. Sequences and their applications: SETA 2004: third international conference, Seoul, Korea, October 24–28, 2004. Revised selected papers, volume 3486 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26084-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26084-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3486;
 http://www.springerlink.
com/openurl.asp?genre=

volume&id=doi:10.1007/
b136167.

Hutter:2007:ALT

[HST07] Marcus Hutter, Rocco A.

Servedio, and Eiji Takimoto, editors. Algorithmic Learning Theory: 18th International Conference, ALT 2007, Sendai, Japan, October 1–4, 2007. Proceedings, volume 4754 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75224-2 (print), 3-540-75225-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75225-7.

Hutter:1999:AFM

[HSTU99] Dieter Hutter, Werner Stephan, Paolo Traverso, and Markus Ullmann, editors. Applied formal methods — FM-Trends 98: International Workshop on Current Trends in Applied Formal Methods, Boppard, Germany, October 7–9, 1998: Proceedings, volume 1641 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66462-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 I59 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1641.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66462-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1641.

Hutter:2010:ALT

[HSVZ10] Marcus Hutter, Frank Stephan, Vladimir Vovk, and Thomas Zeugmann, editors. Algorithmic Learning Theory: 21st International Conference, ALT 2010, Canberra, Australia, October 6–8, 2010. Proceedings, volume 6331 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16107-3 (print), 3-642-16108-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16108-7.

Hameurlain:1997:DES

[HT97] A. Hameurlain and A. Min Tjoa, editors. Database and expert systems applications: 8th International Conference, DEXA’97: Toulouse, France, September 1997: proceedings, volume 1308 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63478-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3D15592 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1308.
htm;   http://www.springerlink.
com/content/978-3-540-
63478-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1308.

Herman:2005:SSS

[HT05] Ted Herman and Sébastien Tixeuil, editors. Self stabilizing systems: 7th international symposium, SSS 2005, Barcelona, Spain, October 26–27, 2005: proceedings, volume 3764 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29814-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S54 W77 2005. URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        29814-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3764.

Hespanha:2006:HSC

[HT06a] João Hespanha and Ashish Tiwari, editors. Hybrid Systems: Computation and Control: 9th International Workshop, HSCC 2006, Santa Barbara, CA, USA, March 29–31, 2006. Proceedings, volume 3927 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33170-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33170-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3927.

Horvath:2006:FMS

[HT06b] András Horváth and Miklós Telek, editors. Formal Methods and Stochastic Models for Performance Evaluation: Third European Performance Engineering Workshop, EPEW 2006, Budapest, Hungary, June 21–22, 2006. Proceedings, volume 4054 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35362-3 (print), 3-540-35365-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35365-2.

Hameurlain:2009:DMG

[HT09] Abdelkader Hameurlain and A. Min Tjoa, editors. Data Management in Grid and Peerto-Peer Systems: Second International Conference, Globe 2009 Linz, Austria, September 1– 2, 2009 Proceedings, volume 5697 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03714-3 (print), 3-642-03715-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03715-3.

Hameurlain:2011:DMG

[HT11] Abdelkader Hameurlain and A. Min Tjoa, editors. Data Management in Grid and Peer-to-Peer Systems: 4th International Conference, Globe 2011, Toulouse, France, September 1–2, 2011, Proceedings, volume 6864 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22946-8 (print), 3-642-22947-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22947-3.

Hong:2009:MEF

[HTMC09] Choong Seon Hong, Toshio Tonouchi, Yan Ma, and Chi-Shih Chao, editors. Management Enabling the Future Internet for Changing Business and New Computing Services: 12th Asia-Pacific Network Operations and Management Symposium, APNOMS 2009 Jeju, South Korea, September 23–25, 2009 Proceedings, volume 5787 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04491-3 (print), 3-642-04492-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04492-2.

Han:2002:EDC

[HTW02] Yanbo Han, Stefan Tai, and Dietmar Wikarski, editors. Engineering and Deployment of Cooperative Information Systems: First International Conference, EDCIS 2002, Beijing, China, September 17– 20, 2002. Proceedings, volume 2480 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44222-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 E33 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2480.htm;
http://www.springerlink.
com/content/978-3-540-
44222-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2480.

Hutter:2005:SPC

[HU05] Dieter Hutter and Markus Ullmann, editors. Security in pervasive computing: Second International Conference, SPC 2005, Boppard, Germany, April 6–8, 2005:

Proceedings, volume 3450 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25521-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5915 .I58 2005.

URL http://www.springerlink.
com/content/978-3-540-
25521-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3450;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b135497.

Heiner:2008:CMS

[HU08] Monika Heiner and Adelinde M. Uhrmacher, editors. Computational Methods in Systems Biology: 6th International Conference CMSB 2008, Rostock, Germany, October 12– 15, 2008. Proceedings, volume 5307 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88561-7 (print), 3-540-88562-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88562-7.

Hu:2009:PLS

[Hu09] Zhenjiang Hu, editor.

Programming Languages and Systems: 7th Asian Symposium, APLAS 2009, Seoul, Korea, December 14–16, 2009. Proceedings, volume 5904 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10671-4 (print), 3-642-10672-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10672-9.

Hughes:1991:FPL

[Hug91] John Hughes, editor.

Functional programming languages and computer architecture: 5th ACM conference, Cambridge, MA, USA, August 26– 30, 1991: proceedings, volume 523 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54396-1 (Berlin), 0-387-54396-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.523. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t0523.htm;
 http://www.springerlink.
com/content/978-3-540-
54396-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=523.

“Fifth conference on Functional Programming Languages and Computer Architecture, held at Harvard University, Cambridge, MA, August 26–30, 1991”–Foreword.

Huget:2003:CMS

[Hug03] Marc-Philippe Huget, editor. Communications in Multiagent Systems: Agent Communication Languages and Conversation Policies, volume 2650 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40385-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 322 pp. LCCN QA76.76.I58 C64 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2650.
htm;   http://www.springerlink.
com/content/978-3-540-
40385-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2650;

http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12446.

Hunt:1994:FVM

[Hun94] Warren A. Hunt. FM8501, a verified microprocessor, volume 795 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-57960-5 (New York), 3-540-57960-5 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 333 pp. LCCN QA73.8.F6 H86 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0795.
htm;   http://www.springerlink.
com/content/978-0-387-
57960-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=795.

Hussmann:1997:FFS

[Hus97] Heinrich Hussmann. Formal foundations for software engineering methods, volume 1322 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63613-7. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 286 pp. LCCN QA76.758 .H87 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1322.htm;
 http://www.springerlink.
com/content/978-3-540-
63613-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1322.

Hussmann:2001:FAS

[Hus01] Heinrich Hussmann, editor. Fundamental approaches to software engineering: 4th International Conference, FASE 2001 held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2001, Genova, Italy, April 2–6, 2001: Proceedings, volume 2029 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41863-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F44 2001; QA267.A1 L43 no.2029. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2029.htm;
 http://www.springerlink.
com/content/978-3-540-
41863-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2029.

Hutchison:1994:MTT

[Hut94] David Hutchison, editor.

Multimedia transport and teleservices: International COST 237 Workshop, Vienna, Austria, November 13–15, 1994: proceedings, volume 882 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58759-4 (Berlin), 0-387-58759-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I65 1994. DM72.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0882.
htm;   http://www.springerlink.
com/content/978-3-540-
58759-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=882.

Hutchison:1996:TMC

[Hut96] David Hutchison, editor. Teleservices and multimedia communications: second International COST 237 Workshop, Copenhagen, Denmark, November 20– 22, 1995: proceedings, volume 1052 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61028-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I65 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1052.htm;
 http://www.springerlink.
com/content/978-3-540-
61028-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1052.

Hu:1998:CAV

[HV98] Alan J. Hu and Moshe Y. Vardi, editors. Computer aided verification: 10th International Conference, CAV’98, Vancouver, BC, Canada, June 28 — July 2, 1998: Proceedings, volume 1427 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64608-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA2667.A1 L43 no.1427. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1427.htm;
 http://www.springerlink.
com/content/978-3-540-
64608-2;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1427.

Hancock:2003:GBR

[HV03] Edwin R. Hancock and Mario Vento, editors. Graph based representations in pattern recognition: 4th IAPR International Workshop, GbRPR 2003, York, UK, June 30–July 2, 2003: Proceedings, volume 2726 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40452-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1650 .G37 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2726.htm;
http://www.springerlink.
com/content/978-3-540-
40452-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2726.

Hermann:2006:LPA

[HV06] Miki Hermann and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 13th International Conference, LPAR 2006, Phnom Penh, Cambodia, November 13–17, 2006. Proceedings, volume 4246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48281-4 (print), 3-540-48282-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48282-6.

HoriaDediu:2012:TPN

[HVT12] Adrian Horia Dediu, Carlos Martín Vide, and Bianca Truthe, editors. Theory and practice of natural computing: first International Conference, TPNC 2012, Tarragona, Spain, October 2–4, 2012. Proceedings, volume 7505 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33859-3, 3-642-33860-7 (e-book). LCCN QA76.9.N37 I58 2012.

URL http://www.springerlink.
com/content/978-3-642-
33859-5;   http://www.
springerlink.com/content/
978-3-642-33860-1.

Helmbold:2001:CLT

[HW01] David Helmbold and Bob Williamson, editors. Computational learning theory: 14th Annual Conference on Computational Learning Theory, COLT 2001 and 5th European Conference on Computational Learning Theory, EuroCOLT 2001, Amsterdam, The Netherlands, July 16–19, 2001: Proceedings, volume 2111 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42343-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2111; Q325.5 .C66 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2111.htm;
 http://www.springerlink.
com/content/978-3-540-
42343-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2111.

Hogrefe:2003:TCS

[HW03] Dieter Hogrefe and Anthony Wiles, editors. Testing of communicating systems: 15th IFIP International Conference, TestCom 2003, Sophia Antipolis, France, May 26–28, 2003: Proceedings, volume 2644 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40123-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5101.A1 I359 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2644.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40123-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2644.

Hung:2005:TAC

[HW05] Dang Hung and Martin Wirsing, editors. Theoretical aspects of computing — ICTAC 2005: second international colloquium, Hanoi, Vietnam, October 17–21, 2005. proceedings, volume 3722 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29107-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29107-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3722.

Hagiya:2006:FLP

[HW06a] Masami Hagiya and Philip Wadler, editors. Functional and Logic Programming: 8th International Symposium, FLOPS 2006, Fuji-Susono, Japan, April 24–26, 2006. Proceedings, volume 3945 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33438-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33438-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3945.

Hong:2006:ADG

[HW06b] Hoon Hong and Dongming Wang, editors. Automated deduction in geometry: 5th international workshop, ADG 2004, Gainesville, FL, USA, September 16– 18, 2004: revised papers, volume 3763 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31332-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA448.D38 I577 2004.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        31332-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3763.

Hochreiter:2007:BRD

[HW07] Sepp Hochreiter and Roland Wagner, editors. Bioinformatics Research and Development: First International Conference, BIRD 2007, Berlin, Germany, March 12–14, 2007. Proceedings, volume 4414 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71232-1 (print), 3-540-71233-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71233-6.

Hu:2008:CCA

[HW08a] Xiaodong Hu and Jie Wang, editors. Computing and Combinatorics: 14th Annual International Conference, COCOON 2008 Dalian, China, June 27–29, 2008 Proceedings, volume 5092 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69732-2 (print), 3-540-69733-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69733-6.

Hudak:2008:PAD

[HW08b] Paul Hudak and David S.

Warren, editors. Practical Aspects of Declarative Languages: 10th International Symposium, PADL 2008, San Francisco, CA, USA, January 7–8, 2008. Proceedings, volume 4902 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77441-6 (print), 3-540-77442-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77442-6.

Hill:2009:LPI

[HW09] Patricia M. Hill and David S. Warren, editors. Logic Programming: 25th International Conference, ICLP 2009, Pasadena, CA, USA, July 14–17, 2009. Proceedings, volume 5649 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02845-4 (print), 3-642-02846-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02846-5.

Heng:2010:CNS

[HWG10] Swee-Huay Heng, Rebecca N. Wright, and Bok-Min Goi, editors. Cryptology and Network Security: 9th International Conference, CANS 2010, Kuala Lumpur, Malaysia, December 12– 14, 2010. Proceedings, volume 6467 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17618-6 (print), 3-642-17619-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17619-7.

Huang:2008:AICb

[HWLJ08a] De-Shuang Huang, Donald C. Wunsch II, Daniel S. Levine, and Kang-Hyun Jo, editors. Advanced Intelligent Computing Theories and Applications. With Aspects of Artificial Intelligence: 4th International Conference on Intelligent Computing, ICIC 2008 Shanghai, China, September 15–18, 2008 Proceedings, volume 5227 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85983-7 (print), 3-540-85984-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85984-0.

Huang:2008:AICa

[HWLJ08b] De-Shuang Huang, Donald C. Wunsch II, Daniel S. Levine, and Kang-Hyun Jo, editors. Advanced Intelligent Computing Theories and Applications. With Aspects of Theoretical and Methodological Issues: 4th International Conference on Intelligent Computing, ICIC 2008 Shanghai, China, September 15–18, 2008 Proceedings, volume 5226 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87440-2 (print), 3-540-87442-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87442-3.

Hannebauer:2001:BRS

[HWP01] Markus Hannebauer, Jan Wendler, and Enrico Pagello, editors. Balancing reactivity and social deliberation in multiagent systems: from RoboCup to real-world applications: A workshop held at the 14th European Conference on Artificial Intelligence (ECAI 2000), Berlin, Germany, August 20–25, 2000, volume 2103 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42327-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 W67 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2103.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-

42327-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2103.

Hancock:2010:SSS

[HWW^{+}10] Edwin R. Hancock, Richard C. Wilson, Terry Windeatt, Ilkay Ulusoy, and Francisco Escolano, editors. Structural, Syntactic, and Statistical Pattern Recognition: Joint IAPR International Workshop, SSPR&SPR 2010, Cesme, Izmir, Turkey, August 18–20, 2010. Proceedings, volume 6218 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14979-0 (print), 3-642-14980-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14980-1.

Huang:2008:AMI

[HXC^{+}08] Yueh-Min Ray Huang, Changsheng Xu, Kuo-Sheng Cheng, Jar-Ferr Kevin Yang, M. N. S. Swamy, Shipeng Li, and Jen-Wen Ding, editors. Advances in Multimedia Information Processing — PCM 2008: 9th Pacific Rim Conference on Multimedia, Tainan, Taiwan, December 9–13, 2008. Proceedings, volume 5353 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89795-X (print), 3-540-89796-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89796-5.

Horvath:2003:ILP

[HY03] Tamás Horváth and Akihiro Yamamoto, editors. Inductive Logic Programming: 13th International Conference, ILP 2003, Szeged, Hungary, September 29– October 1, 2003: Proceedings, volume 2835 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20144-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2835.htm;
 http://www.springerlink.
com/content/978-3-540-
20144-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

0302-9743&volume=2835;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13700.

Hanaoka:2012:AIC

[HY12] Goichiro Hanaoka and Toshihiro Yamauchi, editors. Advances in Information and Computer Security: 7th International Workshop on Security, IWSEC 2012, Fukuoka, Japan, November 7–9, 2012. Proceedings, volume 7631 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34116-0 (print), 3-642-34117-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34117-5.

Hsu:2011:UIC

[HYMZ11] Ching-Hsien Hsu, Laurence T. Yang, Jianhua Ma, and Chunsheng Zhu, editors. Ubiquitous Intelligence and Computing: 8th International Conference, UIC 2011, Banff, Canada, September 2–4, 2011. Proceedings, volume 6905 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23640-5 (print), 3-642-23641-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23641-9.

Huang:2009:ADM

[HYP^{+}09] Ronghuai Huang, Qiang Yang, Jian Pei, João Gama, Xiaofeng Meng, and Xue Li, editors. Advanced Data Mining and Applications: 5th International Conference, ADMA 2009, Beijing, China, August 17–19, 2009. Proceedings, volume 5678 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03347-4 (print), 3-642-03348-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03348-3.

Hsu:2010:AAPa

[HYPY10a] Ching-Hsien Hsu, Laurence T. Yang, Jong Hyuk Park, and Sang-Soo Yeo, editors. Algorithms and Architectures for Parallel Processing: 10th International Conference, ICA3PP 2010, Busan, Korea, May 21– 23, 2010. Proceedings. Part I, volume 6081 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13118-2 (print), 3-642-13119-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13119-6.

Hsu:2010:AAPb

[HYPY10b] Ching-Hsien Hsu, Laurence T. Yang, Jong Hyuk Park, and Sang-Soo Yeo, editors. Algorithms and Architectures for Parallel Processing: 10th International Conference, ICA3PP 2010, Busan, Korea, May 21–23, 2010. Workshops, Part II, volume 6082 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13135-2 (print), 3-642-13136-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        13136-3.

Halbwachs:2005:TAC

[HZ05] Nicolas Halbwachs and Lenore D. Zuck, editors. Tools and algorithms for the construction and analysis of systems: 11th international conference, TACAS 2005, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2005, Edinburgh, UK, April 4–8, 2005: Proceedings, volume 3440 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25333-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T33 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25333-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3440;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b107194.

Holub:2007:IAA

[HŽ07] Jan Holub and Jan Žďárek, editors. Implementation and Application of Automata:

12th International Conference, CIAA 2007, Praque, Czech Republic, July 16–18, 2007, Revised Selected Papers, volume 4783 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76335-X (print), 3-540-76336-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76336-9.

Ho:2008:PTA

[HZ08] Tu-Bao Ho and Zhi-Hua Zhou, editors. PRI-CAI 2008: Trends in Artificial Intelligence: 10th Pacific Rim International Conference on Artificial Intelligence, Hanoi, Vietnam, December 15–19, 2008. Proceedings, volume 5351 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89196-X (print), 3-540-89197-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89197-0.

Hatcliff:2010:FTD

[HZ10] John Hatcliff and Elena Zucca, editors. Formal Techniques for Distributed Systems: Joint 12th IFIP WG 6.1 International Conference, FMOODS 2010 and 30th IFIP WG 6.1 International Conference, FORTE 2010, Amsterdam, The Netherlands, June 7–9, 2010. Proceedings, volume 6117 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13463-7 (print), 3-642-13464-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13464-7.

Horvath:2007:IAF

[HZB07] Zoltán Horváth, Viktória Zsók, and Andrew Butterfield, editors. Implementation and Application of Functional Languages: 18th International Symposium, IFL 2006, Budapest, Hungary, September 4–6, 2006, Revised Selected Papers, volume 4449 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74129-1 (print), 3-540-74130-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74130-5.

Huang:2010:AICa

[HZBF10] De-Shuang Huang, Zhongming Zhao, Vitoantonio Bevilacqua, and Juan Carlos Figueroa, editors. Advanced Intelligent Computing Theories and Applications: 6th International Conference on Intelligent Computing, ICIC 2010, Changsha, China, August 18–21, 2010. Proceedings, volume 6215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14921-9 (print), 3-642-14922-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14922-1.

Huang:2010:AICb

[HZGZ10] De-Shuang Huang, Xiang Zhang, Carlos Alberto Reyes García, and Lei Zhang, editors. Advanced Intelligent Computing Theories and Applications. With Aspects of Artificial Intelligence: 6th International Conference on Intelligent Computing, ICIC 2010, Changsha, China, August 18–21, 2010. Proceedings, volume 6216 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14931-6 (print), 3-642-14932-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14932-0.

Huang:2005:AIC

[HZH05] De-Shuang Huang, Xiao-Ping Zhang, and Guang-Bin Huang, editors. Advances in Intelligent Computing: International Conference on Intelligent Computing, ICIC 2005, Hefei, China, August 23–26, 2005: Proceedings, volume 3644–3645 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28226-2 (vol. 1), 3-540-28227-0 (vol. 2). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q342 .I568 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3644;
http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3645;
http://www.springerlink.
com/content/978-3-540-
28226-6.

Hartmann:2008:WIS

[HZK08] Sven Hartmann, Xiaofang Zhou, and Markus Kirchberg, editors. Web Information Systems Engineering — WISE 2008 Workshops: WISE 2008 International Workshops, Auckland, New Zealand, September 1–4, 2008. Proceedings, volume 5176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85199-2 (print), 3-540-85200-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85200-1.

Huang:2012:NIPa

[HZLL12a] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part I, volume 7663 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34474-7 (print), 3-642-34475-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34475-6.

Huang:2012:NIPb

[HZLL12b] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part II, volume 7664 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34480-1 (print), 3-642-34481-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34481-7.

Huang:2012:NIPc

[HZLL12c] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part III, volume 7665 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34486-0 (print), 3-642-34487-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34487-9.

Huang:2012:NIPd

[HZLL12d] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part IV, volume 7666 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34477-1 (print), 3-642-34478-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34478-7.

Huang:2012:NIPe

[HZLL12e] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part V, volume 7667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34499-2 (print), 3-642-34500-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34500-5.

Ikeda:2006:ITS

[IAC06] Mitsuru Ikeda, Kevin D. Ashley, and Tak-Wai Chan, editors. Intelligent Tutoring Systems: 8th International Conference, ITS 2006, Jhongli, Taiwan, June 26–30, 2006. Proceedings, volume 4053 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35159-0 (print), 3-540-35160-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35160-3.

Ip:2007:AMI

[IAL^{+}07] Horace H.-S. Ip, Oscar C. Au, Howard Leung, Ming-Ting Sun, Wei-Ying Ma, and Shi-Min Hu, editors. Advances in Multimedia Information Processing — PCM 2007: 8th Pacific Rim Conference on Multimedia, Hong Kong, China, December 11– 14, 2007. Proceedings, volume 4810 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77254-5 (print), 3-540-77255-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77255-2.

Ideker:2007:SBC

[IB07] Trey Ideker and Vineet Bafna, editors. Systems Biology and Computational Proteomics: Joint RECOMB 2006 Satellite Workshops on Systems Biology and on Computational Proteomics, San Diego, CA, USA, December 1–3, 2006, Revised Selected Papers, volume 4532 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73059-1 (print), 3-540-73060-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73060-6.

Ibaraki:1992:ACT

[Iba92] Toshihide Ibaraki, editor. Algorithms and computation: Third International Symposium, ISAAC ’92, Nagoya, Japan, December 16–18, 1992: proceedings, volume 650 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-56279-6 (Berlin), 0-387-56279-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I73 1992. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t0650.htm;
http://www.springerlink.
com/content/978-3-540-
56279-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=650.

Ibrahim:1995:SEE

[Ibr95] Rosalind L. Ibrahim, editor. Software engineering education: 8th SEI CSEE Conference, New Orleans, LA, USA, March 29–April 1, 1995: proceedings, volume 895 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-58951-1 (Berlin), 0-387-58951-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S45 1995. DM94.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0895.
htm;   http://www.springerlink.
com/content/978-3-540-
58951-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=895.

“8th Conference on Software Engineering Education (CSEE)”–Pref.

Ichikawa:2007:UCS

[ICSY07] Haruhisa Ichikawa, We-Duke Cho, Ichiro Satoh, and Hee Yong Youn, editors. Ubiquitous Computing Systems: 4th International Symposium, UCS 2007, Tokyo, Japan, November 25– 28, 2007. Proceedings, volume 4836 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76771-1 (print), 3-540-76772-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76772-5.

Ibarra:2003:IAA

[ID03] Oscar H. Ibarra and Zhe Danga, editors. Implementation and application of automata: 8th International Conference, CIAA 2003, Santa Barbara, CA, USA, July 16–18, 2003: Proceedings, volume 2759 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40561-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .C524 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2759.htm;

http://www.springerlink.
com/content/978-3-540-
40561-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2759.

Ibarra:2006:DLT

[ID06] Oscar H. Ibarra and Zhe Dang, editors. Developments in Language Theory: 10th International Conference, DLT 2006, Santa Barbara, CA, USA, June 26–29, 2006. Proceedings, volume 4036 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35428-X (print), 3-540-35430-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35430-7.

IJsselsteijn:2006:PTF

[IdKM06]  Wijnand  IJsselsteijn,

Yvonne de Kort, and Cees Midden, editors. Persuasive Technology: First International Conference on Persuasive Technology for Human Well-Being, PERSUA-SIVE 2006, Eindhoven, The Netherlands, May 2006, Proceedings, volume 3962 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34291-5 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34291-5.

Ishikawa:2008:NIPa

[IDMY08a] Masumi Ishikawa, Kenji Doya, Hiroyuki Miyamoto, and Takeshi Yamakawa, editors. Neural Information Processing: 14th International Conference, ICONIP 2007, Kitakyushu, Japan, November 13–16, 2007, Revised Selected Papers, Part I, volume 4984 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69154-5 (print), 3-540-69158-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69158-7.

Ishikawa:2008:NIPb

[IDMY08b] Masumi Ishikawa, Kenji Doya, Hiroyuki Miyamoto, and Takeshi Yamakawa, editors. Neural Information Processing: 14th International Conference, ICONIP 2007, Kitakyushu, Japan, November 13–16, 2007, Revised Selected Papers, Part II, volume 4985 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69159-6 (print), 3-540-69162-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69162-4.

Ioannides:2010:DHT

[IFGH10] Marinos Ioannides, Dieter Fellner, Andreas Georgopoulos, and Diofantos G. Hadjimitsis, editors. Digital Heritage: Third International Conference, EuroMed 2010, Lemessos, Cyprus, November 8– 13, 2010. Proceedings, volume 6436 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16872-8 (print), 3-642-16873-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16873-4.

Ioannides:2012:PCH

[IFL^{+}12] Marinos Ioannides, Dieter Fritsch, Johanna Leissner, Rob Davies, Fabio Remondino, and Rossella Caffo, editors. Progress in Cultural Heritage Preservation: 4th International Conference, EuroMed 2012, Limassol, Cyprus, October 29 — November 3, 2012. Proceedings, volume 7616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34233-7 (print), 3-642-34234-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34234-9.

Ishida:2007:ICF

[IFV07] Toru Ishida, Susan R. Fussell, and Piek T. J. M. Vossen, editors. Intercultural Collaboration: First International Workshop, IWIC 2007 Kyoto, Japan, January 25–26, 2007 Invited and Selected Papers, volume 4568 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73999-8 (print), 3-540-74000-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74000-1.

Ishida:2005:MMA

[IGN05] Toru Ishida, Les Gasser, and Hideyuki Nakashima, editors. Massively Multi-Agent Systems I: First International Workshop, MMAS 2004, Kyoto, Japan, December 10– 11, 2004, Revised Selected and Invited Papers, volume 3446 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26974-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M62 2004.

URL http://www.springerlink.
com/content/978-3-540-
26974-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3446.

Ito:1990:PLL

[IH90] T. Ito and Robert H.

Halstead, editors. Parallel lisp: languages and systems: US/Japan Workshop on Parallel Lisp, Sendai, Japan, June 5–8, 1989, proceedings, volume 441 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52782-6, 3-540-52782-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.L23 U7 1989. DM52.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0441.htm;
 http://www.springerlink.
com/content/978-0-387-
52782-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=441.

Ito:1996:PSL

[IHQ96] T. (Takayasu) Ito, Robert H. Halstead, and Christian Queinnec, editors. Parallel symbolic languages and systems: International Workshop PSLS’95, Beaune, France, October 2–4, 1995: proceedings, volume 1068 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61143-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I576 1995. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t1068.htm;
http://www.springerlink.
com/content/978-3-540-
61143-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1068.

Ishikawa:2008:AWN

[IHX^{+}08] Yoshiharu Ishikawa, Jing He, Guandong Xu, Yong Shi, Guangyan Huang, Chaoyi Pang, Qing Zhang, and Guoren Wang, editors. Advanced Web and NetworkTechnologies, and Applications: APWeb 2008 International Workshops: BIDM, IWHDM, and DeWeb Shenyang, China, April 26–28, 2008. Revised Selected Papers, volume 4977 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89375-X (print), 3-540-89376-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89376-9.

Ishida:2000:DCT

[II00] Toru Ishida and Katherine Isbister, editors. Digital cities: technologies, experiences, and future perspectives, volume 1765 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67265-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.83. D54 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1765.htm;
 http://www.springerlink.
com/content/978-3-540-
67265-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1765.

Revised selected papers presented at the Kyoto Meeting on Digital Cities, held September 1999.

Inverardi:2006:SEE

[IJ06] Paola Inverardi and Mehdi Jazayeri, editors. Software Engineering Education in the Modern Age: Software Education and Training Sessions at the International Conference on Software Engineering, ICSE 2005, St. Louis, MO, USA, May 15– 21, 2005, Revised Lectures, volume 4309 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68203-1 (print), 3-540-68204-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68204-2.

Isahara:2012:ANL

[IK12] Hitoshi Isahara and Kyoko Kanzaki, editors. Advances in Natural Language Processing: 8th International Conference on NLP, JapTAL 2012, Kanazawa, Japan, October 22–24, 2012. Proceedings, volume 7614 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33982-4 (print), 3-642-33983-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33983-7.

Ito:2008:CGG

[IKKU08] Hiro Ito, Mikio Kano, Naoki Katoh, and Yushi Uno, editors. Computational Geometry and Graph Theory: International Conference, KyotoCGGT 2007, Kyoto, Japan, June 11–15, 2007. Revised Selected Papers, volume 4535 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89549-3 (print), 3-540-89550-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89550-3.

Ibaraki:2003:ACI

[IKO03] Toshihide Ibaraki, Naoki Katoh, and Hirotaka Ono, editors. Algorithms and Computation: 14th International Symposium, ISAAC 2003, Kyoto, Japan, December 15–17, 2003: Proceedings, volume 2906 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20695-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .C684 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2906.htm;
 http://www.springerlink.
com/content/978-3-540-
20695-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2906;
 http://www.springerlink.
com/openurl.asp?genre=

volume&id=doi:10.1007/
b94771.

Ibrahim:2000:DES

[IKR00] Mohamed Ibrahim, Josef Küng, and Norman Revell, editors. Database and Expert Systems Applications: 11th International Conference, DEXA 2000, London, UK, September 4–8, 2000. Proceedings, volume 1873 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67978-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1873. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1873.htm;
http://www.springerlink.
com/content/978-3-540-
67978-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1873.

Ioannidis:2005:ACN

[IKY05] John Ioannidis, Angelos Keromytis, and Moti Yung, editors. Applied Cryptography and Network Security: Third International Conference, ACNS 2005, New York, NY, USA, June 7–10, 2005, proceedings, volume 3531 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26223-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5102.94 .A28 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26223-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3531;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b137093.

Insana:2001:IPM

[IL01] Michael F. Insana and Richard M. Leahy, editors. Information processing in medical imaging: 17th International Conference, IPMI 2001, Davis, CA, USA, June 18–22, 2001: Proceedings, volume 2082 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42245-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2082.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2082.

htm;   http://www.springerlink.
com/content/978-3-540-
42245-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2082.

Iivari:1995:AIS

[ILR95] Juhani Iivari, Kalle Lyytinen, and Matti Rossi, editors. Advanced information systems engineering: 7th International Conference, CAiSE ’95, Jyväskylä, Finland, June 12–16, 1995: proceedings, volume 932 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59498-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.932.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0932.
htm;   http://www.springerlink.
com/content/978-3-540-
59498-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=932.

Ito:1991:TAC

[IM91] T. Ito and Albert R.

Meyer, editors. Theoretical aspects of computer software: International Conference, TACS’91, Sendai, Japan, September 24–27, 1991: proceedings, volume 526 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54415-1 (Berlin), 3-540-54415-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .T33 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0526.htm;
 http://www.springerlink.
com/content/978-0-387-
54415-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=526.

Icard:2010:IEL

[IM10] Thomas Icard and Reinhard Muskens, editors. Interfaces: Explorations in Logic, Language and Computation: ESSLLI 2008 and ESS-LLI 2009 Student Sessions. Selected Papers, volume 6211 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14728-3 (print), 3-642-14729-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
14729-6.

Imam:1999:MAI

[Ima99] Ibrahim Imam, editor. Multiple approaches to intelligent systems: 12th International Conference on Industrial and Engineering Applications of Artificial Intelligence and Expert Systems, IEA/AIE-99, Cairo, Egypt, May 31– June 3, 1999: Proceedings, volume 1611 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66076-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I52 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1611.htm;
http://www.springerlink.
com/content/978-3-540-
66076-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1611.

Ijspeert:2006:BIA

[IMK06] Auke Jan Ijspeert, Toshimitsu Masuzawa, and Shinji Kusumoto, editors. Biologically Inspired Approaches to Advanced Information Technology: Second International Workshop, BIOADIT 2006, Osaka, Japan, January 26–27, 2006, volume 3853 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31253-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31253-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3853.

Ijspeert:2004:BIA

[IMW04] Auke Jan Ijspeert, Masayuki Murata, and Naoki Wakamiya, editors. Biologically Inspired Approaches to Advanced Information Technology: First International Workshop, IBIOADT 2004, Lausanne, Switzerland, January 29–30, 2004, revised Selected Papers, volume 3141 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23339-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t3141.htm;
http://www.springerlink.
com/content/978-3-540-
23339-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3141;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101281.

Iocchi:2009:RRS

[IMWZ09] Luca Iocchi, Hitoshi Matsubara, Alfredo Weitzenfeld, and Changjiu Zhou, editors. RoboCup 2008: Robot Soccer World Cup XII, volume 5399 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02920-5 (print), 3-642-02921-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02921-9.

Indulska:2007:UIC

[IMY^{+}07] Jadwiga Indulska, Jianhua Ma, Laurence T. Yang, Theo Ungerer, and Jiannong Cao, editors. Ubiquitous Intelligence and Computing: 4th International Conference, UIC 2007, Hong Kong, China, July 11–13, 2007.

Proceedings, volume 4611 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73548-8 (print), 3-540-73549-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73549-6.

Iwata:2011:AIC

[IN11] Tetsu Iwata and Masakatsu Nishigaki, editors. Advances in Information and Computer Security: 6th International Workshop, IWSEC 2011, Tokyo, Japan, November 8–10, 2011. Proceedings, volume 7038 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25140-4 (print), 3-642-25141-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25141-2.

Ioannidis:2007:ADI

[INR07] Yannis Ioannidis, Boris Novikov, and Boris Rachev, editors. Advances in Databases and Information Systems: 11th East European Conference, AD-BIS 2007, Varna, Bulgaria, September 29– October 3, 2007. Proceedings, volume 4690 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75184-X (print), 3-540-75185-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75185-4.

Nakano:2008:WAC

[iNR08] Shin ichi Nakano and Md. Saidur Rahman, editors. WALCOM: Algorithms and Computation: Second International Workshop, WALCOM 2008, Dhaka, Bangladesh, February 7– 8, 2008. Proceedings, volume 4921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77890-X (print), 3-540-77891-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        77891-2.

Indulska:2008:PCI

[IPRO08] Jadwiga Indulska, Donald J. Patterson, Tom Rodden, and Max Ott, editors. Pervasive Computing: 6th International Conference, Pervasive 2008 Sydney, Australia, May 19–22, 2008 Proceedings, volume 5013 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79575-8 (print), 3-540-79576-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79576-6.

Iida:2004:EAI

[IPSK04] Fumiya Iida, Rolf Pfeifer, Luc Steels, and Yasuo Kuniyoshi, editors. Embodied Artificial Intelligence: International Seminar, Dagstuhl Castle, Germany, July 7–11, 2003: Revised Selected Papers, volume 3139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22484-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q335 .E43 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3139.htm;
http://www.springerlink.
com/content/978-3-540-
22484-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3139;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99075.

Indulska:2007:DAI

[IR07] Jadwiga Indulska and Kerry Raymond, editors. Distributed Applications and Interoperable Systems: 7th IFIP WG 6.1 International Conference, DAIS 2007, Paphos, Cyprus, June 6–8, 2007. Proceedings, volume 4531 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72881-3 (print), 3-540-72883-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72883-2.

Ibarra:2008:IAA

[IR08] Oscar H. Ibarra and Bala Ravikumar, editors. Implementation and Applications of Automata: 13th International Conference, CIAA 2008, San Francisco, California, USA, July 21–24, 2008. Proceedings, volume 5148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70843-X (print), 3-540-70844-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70844-5.

Imai:1993:ACA

[IRM93] Hideki Imai, Ronald L. Rivest, and Tsutomu Matsumoto, editors. Advances in cryptology, ASIACRYPT ’91: International Conference on the Theory and Application of Cryptology, Fujiyoshida, Japan, November 11–14, 1991: proceedings, volume 739 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-57332-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I555 1991. DM96.00. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t0739.htm;
http://www.springerlink.
com/content/978-0-387-
57332-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=739.

Ip:1998:MIA

[IS98] Horace H. S. Ip and Arnoldus Wilhelmus Maria Smeulders, editors. Multimedia information analysis and retrieval: IAPR International Workshop MINAR’98, Hong Kong, China, August 1998: Proceedings, volume 1464 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64826-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I318 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1464.htm;
http://www.springerlink.
com/content/978-3-540-
64826-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1464.

Ishizuka:2002:PTA

[IS02] Mitsuru Ishizuka and Abdul Sattar, editors. PRICAI 2002: Trends in Artificial Intelligence:

7th Pacific Rim International Conference on Artificial Intelligence, Tokyo, Japan, August 18–22, 2002. Proceedings, volume 2417 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44038-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P33 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2417.htm;
 http://www.springerlink.
com/content/978-3-540-
44038-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2417.

Also available via the World Wide Web.

Issarny:2008:MAI

[IS08] Valérie Issarny and Richard Schantz, editors. Middleware 2008: ACM/IFIP/ USENIX 9th International Middleware Conference Leuven, Belgium, December 1–5, 2008 Proceedings, volume 5346 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89855-7 (print), 3-540-89856-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89856-6.

Iliopoulos:2011:CAIa

[IS11a] Costas S. Iliopoulos and William F. Smyth, editors. Combinatorial Algorithms: 21st International Workshop, IWOCA 2010, London, UK, July 26–28, 2010, Revised Selected Papers, volume 6460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19221-1 (print), 3-642-19222-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19222-7.

Iliopoulos:2011:CAIb

[IS11b] Costas S. Iliopoulos and William F. Smyth, editors. Combinatorial Algorithms: 22nd International Workshop, IWOCA 2011, Victoria, BC, Canada, July 20–22, 2011, Revised Selected Papers, volume 7056 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25010-6 (print), 3-642-25011-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25011-8.

Isokoski:2012:HPDa

[IS12a] Poika Isokoski and Jukka Springare, editors. Haptics: Perception, Devices, Mobility, and Communication: International Conference, EuroHaptics 2012, Tampere, Finland, June 13– 15, 2012. Proceedings, Part I, volume 7282 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31400-7 (print), 3-642-31401-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31401-8.

Isokoski:2012:HPDb

[IS12b] Poika Isokoski and Jukka Springare, editors. Haptics: Perception, Devices, Mobility, and Communication: International Conference, EuroHaptics 2012, Tampere, Finland, June 13– 15, 2012 Proceedings, Part II, volume 7283 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31403-1 (print), 3-642-31404-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31404-9.

Iwamura:2012:CBD

[IS12c] Masakazu Iwamura and Faisal Shafait, editors. Camera-Based Document Analysis and Recognition: 4th International Workshop, CB-DAR 2011, Beijing, China, September 22, 2011, Revised Selected Papers, volume 7139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29363-8 (print), 3-642-29364-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29364-1.

Ishida:1994:PDM

[Ish94]  Toru Ishida.  Paral-

lel, distributed, and multiagent production systems, volume 878 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58698-9 (Berlin), 0-387-58698-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 166 pp. LCCN QA7.58 .I84 1994. DM46.00.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0878.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        58698-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=878.

Ishikawa:1997:SCO

[Ish97] Yutaka Ishikawa, editor. Scientific computing in object-oriented parallel environments: First International Conference, ISCOPE 97, Marina del Rey, California, December 8–11, 1997, proceedings, volume 1343 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63827-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q183.9.I78 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1343.
htm;   http://www.springerlink.
com/content/978-3-540-
63827-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1343.

Ishida:1999:CCS

[Ish99a] Toru Ishida, editor.

Community computing and support systems: social interaction in networked communities, volume 1519 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65475-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.83.C656 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1519.htm;
http://www.springerlink.
com/content/978-3-540-
65475-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1519.

Ishida:1999:MPF

[Ish99b] Toru Ishida, editor.

Multiagent platforms:

First Pacific Rim International Workshop on Multi-Agents, PRIMA ’98, Singapore, November 23, 1998: proceeedings, volume 1599 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65967-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 P75 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1599.
htm;   http://www.springerlink.
com/content/978-3-540-
65967-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1599.

Ishida:2010:CCC

[Ish10] Toru Ishida, editor. Culture and Computing: Computing and Communication for Crosscultural Interaction, volume 6259 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17183-4 (print), 3-642-17184-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17184-0.

Ishai:2011:TCT

[Ish11] Yuval Ishai, editor. Theory of Cryptography: 8th Theory of Cryptography Conference, TCC 2011, Providence, RI, USA, March 28–30, 2011. Proceedings, volume 6597 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19570-9 (print), 3-642-19571-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19571-6.

Ioannidis:2006:ADT

[ISS^{+}06] Yannis Ioannidis, Marc H.

Scholl, Joachim W. Schmidt, Florian Matthes, Mike Hatzopoulos, Klemens Boehm, Alfons Kemper, Torsten Grust, and Christian Boehm, editors. Advances in Database Technology — EDBT 2006: 10th International Conference on Extending Database Technology, Munich, Germany, 26–31 March, 2006, volume 3896 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32960-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32960-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3896.

Inoue:2007:CLM

[IST07] Katsumi Inoue, Ken Satoh, and Francesca Toni, editors. Computational Logic in Multi-Agent Systems: 7th International Workshop, CLIMA VII, Hakodate, Japan, May 8–9, 2006, Revised Selected and Invited Papers, volume 4371 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69618-0 (print), 3-540-69619-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69619-3.

Ito:2003:DLT

[IT03] Masami Ito and Masafumi Toyama, editors. Developments in language theory: 6th International Conference, DLT 2002, Kyoto, Japan, September 18–21, 2002: Revised Papers, volume 2450 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40431-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .C63 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2450.htm;
 http://www.springerlink.
com/content/978-3-540-
40431-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2450.

Iglesias:2006:MSI

[IT06] Andrés Iglesias and Nobuki Takayama, editors. Mathematical Software — ICMS 2006: Second International Congress on Mathematical Software, Castro Urdiales, Spain, September 1– 3, 2006. Proceedings, volume 4151 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38084-1 (print), 3-540-38086-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        38086-3.

Ito:2008:DLT

[IT08] Masami Ito and Masafumi Toyama, editors. Developments in Language Theory: 12th International Conference, DLT 2008, Kyoto, Japan, September 16–19, 2008. Proceedings, volume 5257 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85779-6 (print), 3-540-85780-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85780-8.

Istrail:2004:CMS

[IWC04] Sorin Istrail, Michael Waterman, and Andrew Clark, editors. Computational Methods for SNPs and Haplotype Inference: DIMACS/ RECOMB satellite workshop, Piscataway, NJ, USA, November 21–22, 2002: Revised Papers, volume 2983 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21249-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH438.4.M33 C65 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2983.
htm;   http://www.springerlink.
com/content/978-3-540-
21249-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2983;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96286.

Ito:1995:TPP

[IY95] T. Ito and Akinori Yonezawa, editors. Theory and practice of parallel programming: International Workshop TPPP ’94, Sendai, Japan, November 7–9, 1994: proceedings, volume 907 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 0-387-59172-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .T48 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0907.htm;
 http://www.springerlink.

                                        com/content/978-0-387-
                                        59172-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=907.

Ibarra:2006:IAA

[IY06] Oscar H. Ibarra and Hsu-Chun Yen, editors. Implementation and Application of Automata: 11th International Conference, CIAA 2006, Taipei, Taiwan, August 21–23, 2006. Proceedings, volume 4094 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37213-X (print), 3-540-37214-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37214-1.

Imai:1998:PKC

[IZ98] Hideki Imai and Yuliang Zheng, editors. Public key cryptography: First International Workshop on Practice and Theory in Public Key Cryptography, PKC ’98, Pacifico Yokohama, Japan, February 5–6, 1998: Proceedings, volume 1431 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64693-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I567 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1431.htm;
 http://www.springerlink.
com/content/978-3-540-
64693-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1431.

Imai:1999:PKC

[IZ99] Hideki Imai and Yuliang Zheng, editors. Public key cryptography: Second International Workshop on Practice and Theory in Public Key Cryptography, PKC’99, Kamakura, Japan, March 1–3, 1999: Proceedings, volume 1560 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65644-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I567 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1560.
htm;   http://www.springerlink.
com/content/978-3-540-
65644-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1560.

Imai:2000:PKC

[IZ00] Hideki Imai and Yuliang Zheng, editors. Public key cryptography: Third International Workshop on Practice and Theory in Public Key Cryptosystems, PKC 2000, Melbourne, Victoria, Australia, January 18–20, 2000: Proceedings, volume 1751 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-66967-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1751.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1751.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66967-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1751;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b75033.

Ibarra:2002:CCA

[IZ02] Oscar H. Ibarra and Louxin Zhang, editors. Computing and Combinatorics: 8th Annual International Conference, COCOON 2002, Singapore, August 15– 17, 2002. Proceedings, volume 2387 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43996-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 . C549 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2387.htm;
 http://www.springerlink.
com/content/978-3-540-
43996-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2387.

Iurgel:2009:ISS

[IZP09] Ido A. Iurgel, Nelson Zagalo, and Paolo Petta, editors. Interactive Storytelling: Second Joint International Conference on Interactive Digital Storytelling, ICIDS 2009, Guimarães, Portugal, December 9– 11, 2009. Proceedings, volume 5915 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10642-0 (print), 3-642-10643-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10643-9.

Jacquet:1991:CMA

[Jac91] J.-M. (Jean-Marie) Jacquet. Conclog: a methodological approach to concurrent logic programming, volume 556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54938-2 (Berlin), 0-387-54938-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 781 pp. LCCN QA76.63 .J33 1991. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0556.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        54938-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=556.

Jacobsen:2004:MAI

[Jac04] Hans-Arno Jacobsen, editor. Middleware 2004: ACM/IFIP/USENIX International Middleware Conference, Toronto, Canada, October 18– 22, 2004: proceedings, volume 3231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23428-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3231.

Jacko:2007:HCId

[Jac07a] Julie A. Jacko, editor.

Human-Computer Interaction. HCI Applications and Services: 12th International Conference, HCI International 2007, Beijing, China, July 22– 27, 2007, Proceedings, Part IV, volume 4553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73109-1 (print), 3-540-73111-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73111-5.

Jacko:2007:HCIc

[Jac07b] Julie A. Jacko, editor.

Human-Computer Interaction. HCI Intelligent Multimodal Interaction Environments: 12th International Conference, HCI International 2007, Beijing, China, July 22– 27, 2007, Proceedings, Part III, volume 4552 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73108-3 (print), 3-540-73110-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73110-8.

Jacko:2007:HCIa

[Jac07c] Julie A. Jacko, editor. Human-Computer Interaction. Interaction Design and Usability: 12th International Conference, HCI International 2007, Beijing, China, July 22–27, 2007, Proceedings, Part I, volume 4550 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73104-0 (print), 3-540-73105-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73105-4.

Jacko:2007:HCIb

[Jac07d] Julie A. Jacko, editor. Human-Computer Interaction. Interaction Platforms and Techniques: 12th International Conference, HCI International 2007, Beijing, China, July 22– 27, 2007, Proceedings, Part II, volume 4551 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73106-7 (print), 3-540-73107-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73107-8.

Jacko:2009:HCIc

[Jac09a] Julie A. Jacko, editor.

Human-Computer Interaction. Ambient, Ubiquitous and Intelligent Interaction: 13th International Conference, HCI International 2009, San Diego, CA, USA, July 19–24, 2009, Proceedings, Part III, volume 5612 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02579-X (print), 3-642-02580-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02580-8.

Jacko:2009:HCId

[Jac09b] Julie A. Jacko, editor. Human-Computer Interaction. Interacting in Various Application Domains: 13th International Conference, HCI International 2009, San Diego, CA, USA, July 19–24, 2009, Proceedings, Part IV, volume 5613 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02582-X (print), 3-642-02583-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02583-9.

Jacko:2009:HCIa

[Jac09c] Julie A. Jacko, editor. Human-Computer Interaction. New Trends: 13th International Conference, HCI International 2009, San Diego, CA, USA, July 19–24, 2009, Proceedings, Part I, volume 5610 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02573-0 (print), 3-642-02574-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02574-7.

Jacko:2009:HCIb

[Jac09d] Julie A. Jacko, editor.

Human-Computer Interaction. Novel Interaction Methods and Techniques: 13th International Conference, HCI International 2009, San Diego, CA, USA, July 19–24, 2009, Proceedings, Part II, volume 5611 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02576-5 (print), 3-642-02577-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02577-8.

Jacko:2011:HCIa

[Jac11a] Julie A. Jacko, editor.

Human-Computer Interaction. Design and Development Approaches:

14th International Conference, HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part I, volume 6761 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21601-3 (print), 3-642-21602-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21602-2.

Jacko:2011:HCIb

[Jac11b] Julie A. Jacko, editor. Human-Computer Interaction. Interaction Techniques and Environments: 14th International Conference, HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part II, volume 6762 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21604-8 (print), 3-642-21605-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21605-3.

Jacko:2011:HCIc

[Jac11c] Julie A. Jacko, editor.

Human-Computer Interaction. Towards Mobile and Intelligent Interaction Environments: 14th International Conference, HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part III, volume 6763 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21615-3 (print), 3-642-21616-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21616-9.

Jacko:2011:HCId

[Jac11d] Julie A. Jacko, editor.

Human-Computer Interaction. Users and Applications: 14th International Conference, HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part IV, volume 6764 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21618-8 (print), 3-642-21619-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21619-0.

Jaedicke:2001:NCP

[Jae01] Michael Jaedicke. New Concepts for Parallel Object-Relational Query Processing, volume 2169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42781-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 161 pp. LCCN QA76.58 .J348 2001; QA267.A1 L43 no.2169.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2169.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        42781-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2169.

Jaffar:1999:PPC

[Jaf99] Joxan Jaffar, editor. Principles and practice of constraint programming— CP’99: 5th International Conference, CP’99, Alexandria, VA, USA, October 11–14, 1999: Proceedings, volume 1713 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66626-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 C7 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1713.htm;
 http://www.springerlink.
com/content/978-3-540-
66626-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1713.

Jahne:1993:STI

[Jah93] Bernd Jahne. Spatiotemporal image processing: theory and scientific applications, volume 751 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57418-2 (Berlin), 0-387-57418-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 208 pp. LCCN TA1637 .J36 1993. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0751.
htm;   http://www.springerlink.
com/content/978-3-540-
57418-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=751.

Jahnichen:1999:CCI

[Jäh99] Stefan Jähnichen, editor. Compiler construction: 8th International Conference, CC’99, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS’99, Amsterdam, The Netherlands, March 22–28, 1999: Proceedings, volume 1575 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65717-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1575.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65717-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1575.

Jantke:1987:AII

[Jan87] K. P. Jantke, editor. Analogical and inductive inference: proceedings, volume 265 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18081-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I5771 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0265.htm;
 http://www.springerlink.
com/content/978-0-387-
18081-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=265.

Janssen:1988:TCA

[Jan88] R. Janssen, editor. Trends in computer algebra: International Symposium, Bad Neuenahr, May 19–21, 1987: proceedings, volume 296 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-18928-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 T741 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0296.htm;
 http://www.springerlink.
com/content/978-0-387-
18928-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=296.

Organized by the Scientific Programs Department of IBM Germany and sponsored by IBM Germany.

Jantke:1989:AII

[Jan89] K. P. Jantke, editor. Analogical and inductive inference: International Workshop AII ’89, Reinhardsbrunn Castle, GDR, October, 1989: proceedings, volume 397 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 3-540-51734-0 (Berlin), 0-387-51734-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I577 1989.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0397.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        51734-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=397.

Jantke:1992:AII

[Jan92] K. P. Jantke, editor. Analogical and inductive inference: International Workshop AII ’92, Dagstuhl Castle, Germany, October 1992: proceedings, volume 642 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-56004-1 (Berlin), 0-387-56004-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I577 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0642.
htm;   http://www.springerlink.
com/content/978-3-540-
56004-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=642.

Jantke:1993:ALT

[Jan93] K. P. Jantke, editor. Algorithmic learning theory: 4th international workshop, ALT ’93, Tokyo, Japan, November 8–10, 1993: proceedings, volume 744 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57370-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A48 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0744.
htm;   http://www.springerlink.
com/content/978-0-387-
57370-0;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=744.

Jayanti:1999:DCI

[Jay99] Prasad Jayanti, editor. Distributed computing: 13th International Symposium, DISC’99, Bratislava, Slovak Republic, September 27– 29, 1999: Proceedings, volume 1693 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66531-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D438 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1693.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66531-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1693.

Jayaraman:2004:PAD

[Jay04] Bharat Jayaraman, editor. Practical Aspects of Declarative Languages: 6th International Symposium, PADL 2004, Dallas, TX, USA, June 18–19, 2004: Proceedings, volume 3057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22253-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .P32 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3057.htm;
 http://www.springerlink.
com/content/978-3-540-
22253-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3057;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98355.

Jarke:1994:ADT

[JBJ94] Matthias Jarke, Janis Bubenko, and Keith Jeffery, editors. Advances in database technology

— EDBT ’94: 4th International Conference on Extending Database Technology, Cambridge, United Kingdom, March 28–31, 1994: proceedings, volume 779 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57818-8 (New York), 3-540-57818-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55844 1994. DM80.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0779.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        57818-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=779.

Jensen:2009:TPNb

[JBK09] Kurt Jensen, Jonathan Billington, and Maciej Koutny, editors. Transactions on Petri Nets and Other Models of Concurrency III, volume 5800 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04854-4 (print), 3-642-04856-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04856-2.

Josang:2012:SIS

[JC12] Audun Jøsang and Bengt Carlsson, editors. Secure IT Systems: 17th Nordic Conference, NordSec 2012, Karlskrona, Sweden, October 31 — November 2, 2012. Proceedings, volume 7617 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34209-4 (print), 3-642-34210-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34210-3.

Jeuring:2012:ICM

[JCC^{+}12] Johan Jeuring, John A.

Campbell, Jacques Carette, Gabriel Dos Reis, Petr Sojka, Makarius Wenzel, and Volker Sorge, editors. Intelligent Computer Mathematics: 11th International Conference, AISC 2012, 19th Symposium, Calculemus 2012, 5th International Workshop, DML 2012, 11th International Conference, MKM 2012, Systems and Projects, Held as Part of CICM 2012, Bremen, Germany, July 8–13, 2012. Proceedings, volume 7362 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31373-6 (print), 3-642-31374-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31374-5.

Jiang:2012:ARA

[JDAW12] He Jiang, Wei Ding, Moonis Ali, and Xindong Wu, editors. Advanced Research in Applied Artificial Intelligence: 25th International Conference on Industrial Engineering and Other Applications of Applied Intelligent Systems, IEA/AIE 2012, Dalian, China, June 9–12, 2012. Proceedings, volume 7345 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31086-9 (print), 3-642-31087-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31087-4.

Jensen:2010:TPN

[JDK10] Kurt Jensen, Susanna Donatelli, and Maciej Koutny, editors. Transactions on Petri Nets and Other Models of Concurrency IV, volume 6550 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-18221-6 (print), 3-642-18222-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18222-8.

Jensen:2012:TPNa

[JDK12] Kurt Jensen, Susanna Donatelli, and Jetty Kleijn, editors. Transactions on Petri Nets and Other Models of Concurrency V, volume 6900 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29071-X (print), 3-642-29072-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29072-5.

Jeusfeld:2011:CMI

[JDL11] Manfred Jeusfeld, Lois Delcambre, and Tok-Wang Ling, editors. Conceptual Modeling — ER 2011: 30th International Conference, ER 2011, Brussels, Belgium, October 31 — November 3, 2011. Proceedings, volume 6998 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24605-2 (print), 3-642-24606-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24606-7.

Joseph:2008:APP

[JDMB08] Sam Joseph, Zoran Despotovic, Gianluca Moro, and Sonia Bergamaschi, editors. Agents and Peer-to-Peer Computing: 5th International Workshop, AP2PC 2006, Hakodate, Japan, May 9, 2006, Revised and Invited Papers, volume 4461 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79704-1 (print), 3-540-79705-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79705-0.

Joseph:2010:APP

[JDMB10] Samuel R. H. Joseph, Zoran Despotovic, Gianluca Moro, and Sonia Bergamaschi, editors. Agents and Peer-to-Peer Computing: 6th International Workshop, AP2PC 2007, Honolulu, Hawaii, USA, May 14–18, 2007, Revised and Selected Papers, volume 5319 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11367-2 (print), 3-642-11368-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11368-0.

Jesshope:2006:ACS

[JE06] Chris Jesshope and Colin Egan, editors. Advances in computer systems architecture: 11th Asia-Pacific conference, ACSAC 2006, Shanghai, China, September 6–8, 2006, proceedings, volume 4186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-40056-7 (softcover). LCCN QA76.9.A73 A28 2006. URL http:

//www.springerlink.
com/content/978-3-540-
40056-1.

Jensen:1992:ATP

[Jen92] K. Jensen, editor. Application and theory of Petri nets 1992: 13th international conference, Sheffield, UK, June 22– 26, 1992: proceedings, volume 616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 0-387-55676-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A66 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0616.htm;
 http://www.springerlink.
com/content/978-0-387-
55676-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=616.

Proceedings of the 13th International Conference on Application and Theory of Petri Nets, organized by the School of Computing and Management Sciences at Sheffield City Polytechnic, England.

Jensen:2001:AST

[Jen01] Christian S. Jensen, editor. Advances in spatial and temporal databases: 7th International Symposium, SSTD 2001, Redondo Beach, CA, USA, July 12–15, 2001: Proceedings, volume 2121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42301-X (paperback).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S6925 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2121.htm;
 http://www.springerlink.
com/content/978-3-540-
42301-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2121.

Jensen:2003:EEA

[Jen03] Klaus Jensen, editor.

Experimental and efficient algorithms: Second International Workshop, WEA 2003, Ascona, Switzerland, May 26–28, 2003: Proceedings, volume 2647 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40205-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W43 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2647.htm;
 http://www.springerlink.
com/content/978-3-540-
40205-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2647.

Also available via the World Wide Web.

Jeuring:1998:MPC

[Jeu98] Johan Jeuring, editor. Mathematics of program construction: 4th International Conference, MPC ’98, Marstrand, Sweden, June 15–17, 1998: Proceedings, volume 1422 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64591-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1422. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1422.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64591-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1422.

Jean-Fran:2008:DSI

[JFBH08] Jean-François Jean-Fran, Michael R. Berthold, and Tamás Horváth, editors. Discovery Science: 11th International Conference, DS 2008, Budapest, Hungary, October 13–16, 2008. Proceedings, volume 5255 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88410-6 (print), 3-540-88411-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88411-8.

Jampel:1996:CS

[JFM96] Michael Jampel, Eugene C. Freuder, and Michael Maher, editors. Over-constrained systems, volume 1106 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61479-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .O83 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1106.htm;
http://www.springerlink.
com/content/978-3-540-
61479-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1106.

Papers from a workshop held Sept. 1995 as part of the First International Conference on Principles and Practice of Constraint Programming, at Cassis, France.

Jiang:2011:GBR

[JFT11] Xiaoyi Jiang, Miquel Ferrer, and Andrea Torsello, editors. Graph-Based Representations in Pattern Recognition: 8th IAPR-TC-15 International Workshop, GbRPR 2011, Münster, Germany, May 18–20, 2011. Proceedings, volume 6658 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20843-6 (print), 3-642-20844-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20844-7.

Jin:2004:NPC

[JGXC04] Hai Jin, Guang R. Gao, Zhiwei Xu, and Hao Chen, editors. Network and Parallel Computing: IFIP International Conference, NPC 2004, Wuhan, China, October 18–20, 2004: Proceedings, volume 3222 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23388-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t3222.htm;
http://www.springerlink.
com/content/978-3-540-
23388-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3222;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100357.

Jha:2010:SCL

[Jha10] Girish Nath Jha, editor.

Sanskrit Computational Linguistics: 4th International Symposium, New Delhi, India, December 10–12, 2010. Proceedings, volume 6465 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17527-9 (print), 3-642-17528-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17528-2.

Jezequel:2002:UUM

[JHC02] Jean-Marc Jézéquel, Heinrich Hussmann, and Stephen Cook, editors. UML 2002 — the Unified Modeling Language: model engineering, concepts, and tools: 5th International Conference, Dresden, Germany, September 30– October 4, 2002: Proceedings, volume 2460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44254-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 U45 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2460.htm;
 http://www.springerlink.
com/content/978-3-540-
44254-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2460.

Also available via the World Wide Web with additional title: UML 2002 — The Unified Modeling Language 5th International Conference, Dresden, Germany, September 30–October 4, 2002: Proceedings.

Jeuring:2003:AFP

[JJ03] Johan Jeuring and Simon L. Peyton Jones, editors. Advanced functional programming: 4th international school, AFP 2002, Oxford, UK, August 19– 24, 2002: revised lectures, volume 2638 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40132-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .A38 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2638.htm;
 http://www.springerlink.
com/content/978-3-540-
40132-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2638.

Jorge:2003:ISD

[JJF03] Joaquim A. Jorge, Nuno Jardim Nunes, and João Falcao e Cunha, editors. Interactive Systems. Design, Specification, and Verification: Design, specification, and verification: 10th International Workshop, DSV-IS 2003, Funchal, Madeira Islands, Portugal, June 11–13, 2003: Revised Papers, volume 2844 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20159-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.I58. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2844.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20159-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2844;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13960.

Jensen:2002:ADT

[JJP^{+}02] Christian S. Jensen, Keith G. Jeffery, Jaroslav Pokorny, Simonas Šaltenis, Elisa Bertino, Klemens Böhm, and Matthias Jarke, editors. Advances in database technology — EDBT 2002: 8th International Conference on Extending Database Technology, Prague, Czech Republic, March 25–27, 2002: Proceedings, volume 2287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43324-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55844 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2287.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        43324-8;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2287.

Jansen:2000:AAC

[JK00] Klaus Jansen and Samir Khuller, editors. Approximation Algorithms for Combinatorial Optimization: Third International Workshop, APPROX 2000, Saarbrücken, Germany, September 5– 8, 2000. Proceedings, volume 1913 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67996-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W67 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1913.htm;
http://www.springerlink.
com/content/978-3-540-
67996-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1913.

Jonassen:2004:ABI

[JK04] Inge Jonassen and Junhyong Kim, editors. Algorithms in Bioinformatics: 4th International Workshop, WABI 2004, Bergen, Norway, September 17–21, 2004: Proceedings, volume 3240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23018-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3240.htm;
 http://www.springerlink.
com/content/978-3-540-
23018-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3240;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100405.

Julisch:2005:IMD

[JK05a] Klaus Julisch and Christopher Kruegel, editors. Intrusion and malware detection and vulnerability assessment: second international conference, DIMVA 2005, Vienna, Austria, July 7–8, 2005, proceedings, volume 3548 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26613-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26613-6;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3548;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b137798.

Junger:2005:IPC

[JK05b] Michael Jünger and Volker Kaibel, editors. Integer programming and combinatorial optimization: 11th international IPCO conference, Berlin, Germany, June 8–10, 2005. Proceedings, volume 3509 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26199-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26199-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3509;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b137143.

Julliand:2006:BFS

[JK06] Jacques Julliand and Olga Kouchnarenko, editors. B 2007: Formal Specification and Development in B: 7th International Conference of B Users, Besançon, France, January 17– 19, 2007. Proceedings, volume 4355 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68760-2 (print), 3-540-68761-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68761-0.

Jeffery:1996:STP

[JKB96] Keith (Keith G.) Jeffery, Jaroslav Král, and Miroslav Bartosek, editors. SOFSEM’96: theory and practice of informatics: 23rd Seminar on Current Trends in Theory and Practice of Informatics, Milovy, Czech Republic, November 23–30, 1996: proceedings, volume 1175 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61994-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751.S62 1996. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t1175.htm;
http://www.springerlink.
com/content/978-3-540-
61994-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1175.

Jeckle:2004:GSE

[JKB04] Mario Jeckle, Ryszard Kowalczyk, and Peter Braun, editors. Grid services engineering and management: first international conference, GSEM 2004, Erfurt, Germany, September 27– 30, 2004: proceedings, volume 3270 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23301-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 G79 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23301-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3270;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100715.

Jarke:2002:KAA

[JKL02] Matthias Jarke, Jana Koehler, and Gerhard Lakemeyer, editors. KI 2002: Advances in Artificial Intelligence: 25th Annual German Conference on AI, KI 2002, Aachen, Germany, September 16–20, 2002. Proceedings, volume 2479 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44185-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G47 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2479.htm;
 http://www.springerlink.
com/content/978-3-540-
44185-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2479.

Jezic:2012:AMA

[JKN^{+}12] Gordan Jezic, Mario Kusek, Ngoc-Thanh Nguyen, Robert J. Howlett, and Lakhmi C. Jain, editors. Agent and Multi-Agent Systems. Technologies and Applications: 6th KES International Conference, KES-AMSTA 2012,Dubrovnik, Croatia, June 25–27, 2012. Proceedings, volume 7327 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30946-1 (print), 3-642-30947-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30947-2.

Jansen:2004:ARC

[JKRR04] Klaus Jansen, Sanjeev Khanna, José D. P. Rolim, and Dana Ron, editors. Approximation, Randomization, and Combinatorial Optimization: 7th International Workshop on Approximation Algorithms for Combinatorial Optimization Problems, AP-PROX 2004 and 8th International Workshop on Randomization and Computation, RANDOM 2004 Cambridge, MA, USA, August 22–24, 2004 Proceedings, volume 3122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22894-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I643 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3122.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        22894-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3122;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99805.

Jantke:1995:ALK

[JL95] K. P. Jantke and Steffen Lange. Algorithmic learning for knowledgebased systems: GOSLER final report, volume 961 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60217-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 510 pp. LCCN QA76.76.E95 A47 1995.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0961.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60217-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=961.

Jiang:1997:CCT

[JL97] Tao Jiang and D. T. Lee, editors. Computing and combinatorics: third annual international conference, COCOON ’97, Shanghai, China, August 20–22, 1997: proceedings, volume 1276 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63357-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5.I535 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1276.
htm;   http://www.springerlink.
com/content/978-3-540-
63357-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1276.

Jennings:2000:IAV

[JL00] Nick Jennings and Yves Lesperance, editors. Intelligent agents VI: agent theories, architectures, and languages: 6th International Workshop (ATAL’99), Orlando, Florida, USA, July 15– 17, 1999: Proceedings, volume 1757 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67200-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1757. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1757.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67200-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1757.

Jajodia:2008:CSE

[JL08] Sushil Jajodia and Javier Lopez, editors. Computer Security — ES-ORICS 2008: 13th European Symposium on Research in Computer Security, Málaga, Spain, October 6–8, 2008. Proceedings, volume 5283 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88312-6 (print), 3-540-88313-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88313-5.

Jones:2011:DHC

[JL11] Cliff B. Jones and John L. Lloyd, editors. Dependable and Historic Computing: Essays Dedicated to Brian Randell on the Occasion of His 75th Birthday, volume 6875 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24540-4 (print), 3-642-24541-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24541-1.

Jazayeri:2000:GPI

[JLM00] Mehdi Jazayeri, R. Loos, and David R. Musser, editors. Generic programming: International Seminar on Generic Programming, Dagstuhl Castle, Germany, April 27– May 1, 1998: Selected Papers, volume 1766 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41090-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6245 .I58 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1766.htm;
 http://www.springerlink.
com/content/978-3-540-
41090-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1766.

Jeschke:2011:IRAa

[JLS11a] Sabina Jeschke, Honghai Liu, and Daniel Schilberg, editors. Intelligent Robotics and Applications: 4th International Conference, ICIRA 2011, Aachen, Germany, December 6–8, 2011, Proceedings, Part I, volume 7101 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25485-3 (print), 3-642-25486-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25486-4.

Jeschke:2011:IRAb

[JLS11b] Sabina Jeschke, Honghai Liu, and Daniel Schilberg, editors. Intelligent Robotics and Applications: 4th International Conference, ICIRA 2011, Aachen, Germany, December 6–8, 2011, Proceedings, Part II, volume 7102 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25488-8 (print), 3-642-25489-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25489-5.

Jantke:2006:FWI

[JLST06] Klaus P. Jantke, Aran Lunzer, Nicolas Spyratos, and Yuzuru Tanaka, editors. Federation over the Web: International Workshop, Dagstuhl Castle, Germany, May 1–6, 2005. Revised Selected Papers, volume 3847 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31018-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31018-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3847.

Jansen:2002:AAC

[JLV02] Klaus Jansen, S. Leonardi, and Vijay V. Vazirani, editors. Approximation Algorithms for Combinatorial Optimization: 5th International Workshop, APPROX 2002, Rome, Italy, September 17–21, 2002. Proceedings, volume 2462 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44186-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A65 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2462.htm;
 http://www.springerlink.
com/content/978-3-540-
44186-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2462.

Jones:2007:FMH

[JLW07a] Cliff B. Jones, Zhiming Liu, and Jim Woodcock, editors. Formal Methods and Hybrid Real-Time Systems: Essays in Honor of Dines Bjørner and Chaochen Zhou on the Occasion of Their 70th Birthdays, volume 4700 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75220-X (print), 3-540-75221-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
75221-9.

Jones:2007:TAC

[JLW07b] Cliff B. Jones, Zhiming Liu, and Jim Woodcock, editors. Theoretical Aspects of Computing — ICTAC 2007: 4th International Colloquium, Macau, China, September 26–28, 2007. Proceedings, volume 4711 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75290-0 (print), 3-540-75292-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75292-9.

James:2003:NHI

[JLY03] Anne James, B. J. Lings, and Muhammad Younas, editors. New horizons in information management: 20th British National Conference on Databases, BNCOD 20, Coventry, UK, July 15– 17, 2003: Proceedings, volume 2712 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40536-4 (paperback).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B75 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2712.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40536-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2712.

Jones:1978:TUT

[JM78] Neil D. Jones and Steven S. Muchnick. TEMPO: a unified treatment of binding time and parameter passing concepts in programming languages, volume 66 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1978. CO-DEN LNCSD9. ISBN 0-387-09085-1. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 118 pp. LCCN QA76.73.T45 J66. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0066.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        09085-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=66.

Jeuring:1995:AFP

[JM95] Johan Jeuring and Erik Meijer, editors. Advanced functional programming: First International Spring School on Advanced Functional Programming Techniques, Bastad, Sweden, May 24–30, 1995: tutorial text, volume 925 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59451-5 (Berlin), 0-387-59451-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .I59 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0925.htm;
 http://www.springerlink.
com/content/978-3-540-
59451-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=925.

Jean-Marie:2002:ACS

[JM02] A. Jean-Marie, editor.

Advances in computing science — ASIAN 2002: Internet computing and modeling, grid computing, peer-to-peer computing, and cluster computing: 7th Asian Computing Science Conference, Hanoi, Vietnam, December 4–6, 2002: Proceedings, volume 2550 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00195-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A85 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2550.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00195-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2550.

Also available via the

World Wide Web.

Johansson:2003:PCI

[JM03] Thomas Johansson and Subhamoy Maitra, editors. Progress in Cryptology—INDOCRYPT 2003: 4th International Conference on Cryptology in India, New Delhi, India, December 8–10, 2003: Proceedings, volume 2904 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20609-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I5535 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2904.

htm;   http://www.springerlink.
com/content/978-3-540-
20609-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2904;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94593.

Jajodia:2005:ISS

[JM05] Sushil Jajodia and Chandan Mazumdar, editors. Information Systems Security: First International Conference, ICISS 2005, Kolkata, India, December 19–21, 2005. Proceedings, volume 3803 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30706-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30706-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3803.

Janowski:2007:DCI

[JM07] Tomasz Janowski and Hrushikesha Mohanty, editors. Distributed Computing and Internet Technology: 4th International Conference, ICD-CIT 2007, Bangalore, India, December 17–20. Proceedings, volume 4882 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77112-3 (print), 3-540-77115-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77115-9.

Janowski:2010:DCI

[JM10] Tomasz Janowski and Hrushikesha Mohanty, editors. Distributed Computing and Internet Technology: 6th International Conference, ICD-CIT 2010, Bhubaneswar, India, February 15– 17, 2010. Proceedings, volume 5966 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11658-2 (print), 3-642-11659-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11659-9.

Jajodia:2011:ISS

[JM11a] Sushil Jajodia and Chandan Mazumdar, editors.

Information Systems Security: 7th International Conference, ICISS 2011, Kolkata, India, December 15–19, 2011, Proceedings, volume 7093 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25559-0 (print), 3-642-25560-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25560-1.

Jha:2011:ISS

[JM11b] Somesh Jha and Anish Mathuria, editors. Information Systems Security: 6th International Conference, ICISS 2010, Gandhinagar, India, December 17–19, 2010. Proceedings, volume 6503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17713-1 (print), 3-642-17714-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17714-9.

Jahne:2007:CMF

[JMBS07] Bernd Jähne, Rudolf Mester, Erhardt Barth, and Hanno Scharr, editors. Complex Motion: First International Workshop, IWCM 2004, Günzburg, Germany, October 12–14, 2004. Revised Papers, volume 3417 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69864-7 (print), 3-540-69866-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69866-1.

Jiang:2010:MMP

[JMC10] Xiaoyi Jiang, Matthew Y. Ma, and Chang Wen Chen, editors. Mobile Multimedia Processing: Fundamentals, Methods, and Applications, volume 5960 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12348-1 (print), 3-642-12349-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12349-8.

Josang:2009:IPI

[JMK09] Audun Jøsang, Torleiv Maseng, and Svein Johan Knapskog, editors. Identity and Privacy in the Internet Age: 14th Nordic Conference on Secure IT Systems, NordSec 2009, Oslo, Norway, 14–16 October 2009. Proceedings, volume 5838 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04765-3 (print), 3-642-04766-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04766-4.

Jones:2009:VMC

[JMO09] Neil D. Jones and Markus Müller-Olm, editors. Verification, Model Checking, and Abstract Interpretation: 10th International Conference, VMCAI 2009, Savannah, GA, USA, January 18–20, 2009. Proceedings, volume 5403 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-93899-0 (print), 3-540-93900-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        93900-9.

Joye:2010:PBC

[JMO10] Marc Joye, Atsuko Miyaji, and Akira Otsuka, editors. Pairing-Based Cryptography — Pairing 2010: 4th International Conference, Yamanaka Hot Spring, Japan, December 2010. Proceedings, volume 6487 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17454-X (print), 3-642-17455-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17455-1.

Joshi:2012:VST

[JMP12] Rajeev Joshi, Peter Müller, and Andreas Podelski, editors. Verified Software: Theories, Tools, Experiments: 4th International Conference, VSTTE 2012, Philadelphia, PA, USA, January 28–29, 2012. Proceedings, volume 7152 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27704-7 (print), 3-642-27705-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27705-4.

Joye:2011:SAI

[JMT11] Marc Joye, Debdeep Mukhopadhyay, and Michael Tunstall, editors. Security Aspects in Information Technology: First International Conference, InfoSecHiCom-Net 2011, Haldia, India, October 19–22, 2011. Proceedings, volume 7011 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24585-4 (print), 3-642-24586-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24586-2.

Junger:2001:CCO

[JN01] Michael Jünger and Denis Naddef, editors.

Computational Combinatorial Optimization: Optimal or Provably Near-Optimal Solutions, volume 2241 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42877-1. ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 303 pp. LCCN QA267.A1 L43 no.2241; T57.7 .C66 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2241.htm;
 http://www.springerlink.
com/content/978-3-540-
42877-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2241.

Janhunen:2010:LAI

[JN10] Tomi Janhunen and Ilkka Niemelä, editors. Logics in Artificial Intelligence: 12th European Conference, JELIA 2010, Helsinki, Finland, September 13–15, 2010. Proceedings, volume 6341 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15674-6 (print), 3-642-15675-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15675-5.

Jurdzinski:2012:FMA

[JN12] Marcin Jurdziński and Dejan Ničković, editors. Formal Modeling and Analysis of Timed Systems: 10th International Conference, FORMATS 2012, London, UK, September 18– 20, 2012. Proceedings, volume 7595 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33364-8 (print), 3-642-33365-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33365-1.

Jedrzejowicz:2011:CCIa

[JNH11a] Piotr Jedrzejowicz, Ngoc Thanh Nguyen, and Kiem Hoang, editors. Computational Collective Intelligence. Technologies and Applications: Third International Conference, ICCCI 2011, Gdynia, Poland, September 21– 23, 2011, Proceedings, Part I, volume 6922 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23934-X (print), 3-642-23935-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23935-9.

Jedrzejowicz:2011:CCIb

[JNH11b] Piotr Jedrzejowicz, Ngoc Thanh Nguyen, and Kiem Hoang, editors. Computational Collective Intelligence. Technologies and Applications: Third International Conference, ICCCI 2011, Gdynia, Poland, September 21– 23, 2011, Proceedings, Part II, volume 6923 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23937-4 (print), 3-642-23938-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23938-0.

Jedrzejowicz:2010:AMAa

[JNHJ10a] Piotr Jedrzejowicz, Ngoc Thanh Nguyen, Robert J. Howlet, and Lakhmi C. Jain, editors. Agent and Multi-Agent Systems: Technologies and Applications: 4th KES International Symposium, KES-AMSTA 2010, Gdynia, Poland, June 23–25, 2010, Proceedings. Part I, volume 6070 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13479-3 (print), 3-642-13480-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13480-7.

Jedrzejowicz:2010:AMAb

[JNHJ10b] Piotr Jedrzejowicz, Ngoc Thanh Nguyen, Robert J. Howlet, and Lakhmi C. Jain, editors. Agent and Multi-Agent Systems: Technologies and Applications: 4th KES International Symposium, KES-AMSTA 2010, Gdynia, Poland, June 23–25, 2010, Proceedings. Part II, volume 6071 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13540-4 (print), 3-642-13541-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13541-5.

Jiang:2010:MICa

[JNPV10a] Tianzi Jiang, Nassir Navab, Josien P. W. Pluim, and Max A. Viergever, editors. Medical Image Computing and Computer-Assisted Intervention — MIC-CAI 2010: 13th International Conference, Beijing, China, September 20–24, 2010, Proceedings, Part I, volume 6361 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15704-1 (print), 3-642-15705-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15705-9.

Jiang:2010:MICb

[JNPV10b] Tianzi Jiang, Nassir Navab, Josien P. W. Pluim, and Max A. Viergever, editors. Medical Image Computing and Computer-Assisted Intervention — MIC-CAI 2010: 13th International Conference, Beijing, China, September 20–24, 2010, Proceedings, Part II, volume 6362 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15744-0 (print), 3-642-15745-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15745-5.

Jiang:2010:MICc

[JNPV10c] Tianzi Jiang, Nassir Navab, Josien P. W. Pluim, and Max A. Viergever, editors. Medical Image Computing and Computer-Assisted Intervention — MIC-CAI 2010: 13th International Conference, Beijing, China, September 20–24, 2010, Proceedings, Part III, volume 6363 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15710-6 (print), 3-642-15711-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15711-0.

Jeannot:2011:EPPa

[JNR11a] Emmanuel Jeannot, Raymond Namyst, and Jean Roman, editors. Euro-Par 2011 Parallel Processing: 17th International Conference, Euro-Par 2011, Bordeaux, France, August 29 — September 2, 2011, Proceedings, Part I, volume 6852 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23399-6 (print), 3-642-23400-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23400-2.

Jeannot:2011:EPPb

[JNR11b] Emmanuel Jeannot, Raymond Namyst, and Jean Roman, editors. Euro-Par 2011 Parallel Processing: 17th International Conference, Euro-Par 2011, Bordeaux, France, August 29 — September 2, 2011, Proceedings, Part II, volume 6853 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23396-1 (print), 3-642-23397-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23397-5.

Jackson:2005:DES

[JNS05a] Mike Jackson, David Nelson, and Sue Stirk, editors. Database: Enterprise, Skills and Innovation: 22nd British National Conference on Databases, BNCOD 22, Sunderland, UK, July 5–7, 2005. Proceedings, volume 3567 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26973-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D53 B75 2005.

URL http://www.springerlink.
com/content/978-3-540-
26973-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3567.

Jermann:2005:GOC

[JNS05b]  Christophe  Jermann,

Arnold Neumaier, and Djamila Sam, editors. Global optimization and constraint satisfaction: second international workshop, COCOS 2003, Lausanne, Switzerland, November 18–21, 2003. Revised selected papers, volume 3478 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26003-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26003-5;   http://www.
springerlink.com/link.
asp?id=u2yf529g4p53;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3478;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136292.

Jarke:1999:AIS

[JO99] Matthias Jarke and A. Oberweis, editors. Advanced information systems engineering: 11th International Conference, CAiSE ’99, Heidelberg, Germany, June 14–18, 1999: Proceedings, volume 1626 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66157-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .N67 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1626.htm;
 http://www.springerlink.
com/content/978-3-540-
66157-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1626.

Jajodia:1998:AMI

[JOD98] Sushil Jajodia, M. Tamer Ozsu, and Asuman Dogac, editors. Advances in multimedia information systems: 4th International Symposium, MIS ’98, Istanbul, Turkey, September 24–26, 1998: Proceedings, volume 1508 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65107-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575.I63 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1508.htm;
 http://www.springerlink.
com/content/978-3-540-
65107-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1508.

Johnson:1997:AMS

[Joh97] Michael Johnson, editor. Algebraic methodology and software technology: 6th international conference, AMAST 97, Sydney, Australia, December 13–17, 1997: proceedings, volume 1349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63888-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758.I5713 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1349.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63888-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1349.

Johnson:2001:ISD

[Joh01] Chris Johnson, editor. Interactive systems: design, specification, and verification: 8th International Workshop, DSV-IS 2001, Glasgow, Scotland, UK, June 13– 15: Revised Papers, volume 2220 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42807-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.I58 D78 2001; QA267.A1 L43 no.2220. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2220.htm;
 http://www.springerlink.
com/content/978-3-540-
42807-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2220.

Johansson:2003:FSE

[Joh03] Thomas Johansson, editor. Fast Software Encryption: 10th International Workshop, FSE 2003, Lund, Sweden, February 24–26, 2003: Revised Papers, volume 2887 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20449-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F77 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2887.
htm;   http://www.springerlink.
com/content/978-3-540-
20449-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2887;

                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b93938.

Jones:1980:SDC

[Jon80] Neil D. Jones, editor. Semantics-directed compiler generation: proceedings of a workshop, Aarhus, Denmark, January, 1980, volume 94 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10250-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .W698 1980. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0094.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        10250-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=94.

Jones:1990:EES

[Jon90] Neil D. Jones, editor. ESOP ’90: 3rd European Symposium on Programming, Copenhagen, Denmark, May 15–18, 1990: proceedings, volume 432 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52592-0 (Berlin), 0-387-52592-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0432.htm;
 http://www.springerlink.
com/content/978-3-540-
52592-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=432.

Jonker:2000:DTI

[Jon00] Willem Jonker, editor.

Databases in telecommunications: International Workshop co-located with VLDB-99, Edinburgh, Scotland, UK, September 6th, 1999: Proceedings, volume 1819 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67667-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D3645 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1819.htm;
 http://www.springerlink.
com/content/978-3-540-
67667-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1819.

Jonker:2001:DTI

[Jon01] Willem Jonker, editor. Databases in telecommunications II: VLDB 2001 International Workshop, Rome, Italy, September 10, 2001: Proceedings, volume 2209 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42623-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D3645 2001; QA267.A1 L43 no.2209.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2209.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        42623-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2209.

Refereed proceedings of the Second International Workshop on Databases in Telecommunications, Held during VLDB 2001 in Rome, Italy in September 2001.

Jonasson:2012:APSa

[Jón12a] Kristján Jónasson, editor. Applied Parallel and Scientific Computing: 10th International Conference, PARA 2010, Reykjavík, Iceland, June 6–9, 2010, Revised Selected Papers, Part I, volume 7133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28150-8 (print), 3-642-28151-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28151-8.

Jonasson:2012:APSb

[Jón12b] Kristján Jónasson, editor. Applied Parallel and Scientific Computing: 10th International Conference, PARA 2010, Reykjavík, Iceland, June 6–9, 2010, Revised Selected Papers, Part II, volume 7134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28144-3 (print), 3-642-28145-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28145-7.

Joseph:1988:FTR

[Jos88] M. Joseph, editor. Formal techniques in realtime and fault-tolerant systems: proceedings of a symposium, Warwick, UK, September 22–23, 1988, volume 331 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-50302-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54 .F661 1988. DM36.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0331.htm;
 http://www.springerlink.
com/content/978-0-387-
50302-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=331.

Papers from the Symposium on Formal Techniques in Real-Time and Fault-Tolerant Systems held at the University of Warwick.

Joseph:2000:FTR

[Jos00] M. Joseph, editor. Formal Techniques in Real-Time and Fault-Tolerant Systems: 6th International Symposium, FTRTFT 2000, Pune, India, September 20–22, 2000. Proceedings, volume 1926 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41055-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1.L43 no.1926. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1926.htm;
 http://www.springerlink.
com/content/978-3-540-
41055-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1926.

Jouannaud:1985:FPL

[Jou85a] Jean-Pierre Jouannaud, editor. Functional programming languages and computer architecture: Nancy, France, September 1985, volume 201 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15975-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.201. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0201.htm;
 http://www.springerlink.
com/content/978-0-387-
15975-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=201.

Jouannaud:1985:RTA

[Jou85b] Jean-Pierre Jouannaud, editor. Rewriting techniques and applications: Dijon, France, May 20– 22, 1985, volume 202 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15976-2 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .R4641 1985. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0202.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        15976-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=202.

Papers presented at the First International Conference on Rewriting Techniques and Applications, held in Dijon, Burgundy, France, May 20– 22, 1985.

Jouannaud:1994:CCL

[Jou94] Jean-Pierre Jouannaud, editor. Constraints in computational logics: first international conference, CCL ’94, Munich, Germany, September 7–9, 1994: proceedings, volume 845 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-58403-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA74.6 .C6515 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0845.htm;
http://www.springerlink.
com/content/978-0-387-
58403-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=845.

Joux:2009:ACE

[Jou09] Antoine Joux, editor.

Advances in cryptology

— EUROCRYPT 2009: 28th Annual International Conference on the Theory and Applications of Cryptographic Techniques, Cologne, Germany, April 26– 30, 2009. Proceedings, volume 5479. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-642-01001-6, 3-642-01000-8. LCCN QA76.9.A25 E97 2009. URL http:

//www.springerlink.
com/content/978-3-642-
01001-9.

Joux:2011:FSE

[Jou11] Antoine Joux, editor.

Fast Software Encryption: 18th International Workshop, FSE 2011, Lyngby, Denmark, February 13–16, 2011, Revised Selected Papers, volume 6733 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21701-X (print), 3-642-21702-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21702-9.

Joye:2003:TCC

[Joy03] Marc Joye, editor. Topics in cryptology, CT-RSA 2003: the Cryptographers’ Track at the RSA Conference 2003, San Francisco, CA, USA, April 13–17, 2003: Proceedings, volume 2612 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00847-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C822 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2612.
htm;   http://www.springerlink.
com/content/978-3-540-
00847-7;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2612.

Jonsson:1994:CCT

[JP94] Bengt Jonsson and Joachim Parrow, editors. CON-CUR ’94, concurrency theory: 5th international conference, Uppsala, Sweden, August 22–25, 1994: proceedings, volume 836 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58329-7 (Berlin), 0-387-58329-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .C66 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0836.htm;
http://www.springerlink.
com/content/978-3-540-
58329-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=836.

Jonsson:1996:FTR

[JP96] Bengt Jonsson and Joachim Parrow, editors. Formal techniques in realtime and fault-tolerant systems: 4th International Symposium, Uppsala, Sweden, September 9–13, 1996: proceedings, volume 1135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61648-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54 .F657 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1135.htm;
 http://www.springerlink.
com/content/978-3-540-
61648-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1135.

Jeusfeld:2003:CMN

[JP03] Manfred A. Jeusfeld and Óscar Pastor, editors. Conceptual Modeling for Novel Application Domains: ER 2003 workshops ECOMO, IWCMQ, AOIS, and XSDM, Chicago, IL, USA, October 13, 2003: Proceedings, volume 2814 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20257-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2814.htm;
 http://www.springerlink.
com/content/978-3-540-

20257-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2814;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13245.

Jensen:2004:TAC

[JP04a] Kurt Jensen and Andreas Podelski, editors. Tools and Algorithms for the Construction and Analysis of Systems: 10th International Conference, TACAS 2004, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2004, Barcelona, Spain, March 29–April 2, 2004: Proceedings, volume 2988 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21299-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T33 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2988.
htm;   http://www.springerlink.
com/content/978-3-540-
21299-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2988;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b96393.

Jonker:2004:SDM

[JP04b] Willem Jonker and Milan Petković, editors. Secure data management: VLDB 2004 Workshop, SDM 2004, Toronto, Canada, August 30, 2004: proceedings, volume 3178 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22983-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D314 S46 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        22983-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3178.

Jacquet:2005:CML

[JP05a] Jean-Marie Jacquet and Gian Pietro Picco, editors. Coordination models and languages: 7th international conference, COORDINATION 2005, Namur, Belgium, April 20–23, 2005: Proceedings, volume 3454 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25630-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .C695 2004. URL http:

//www.springerlink.
com/content/978-3-540-
25630-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3454;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b135676.

Jonker:2005:SDM

[JP05b] Willem Jonker and Milan Petkovic, editors. Secure Data Management: Second VLDB Workshop, SDM 2005, Trondheim, Norway, August 30–September 2, 2005. Proceedings, volume 3674 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28798-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28798-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3674.

Jonker:2006:SDM

[JP06] Willem Jonker and Milan Petković, editors. Secure Data Management: Third VLDB Workshop, SDM 2006, Seoul, Korea, September 10–11, 2006. Proceedings, volume 4165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38984-9 (print), 3-540-38987-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38987-3.

Janssen:2007:CAA

[JP07a] Jeannette Janssen and Paweł Prałat, editors. Combinatorial and Algorithmic Aspects of Networking: 4th Workshop, CAAN 2007, Halifax, Canada, August 14, 2007. Revised Papers, volume 4852 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77293-6 (print), 3-540-77294-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77294-1.

Jonker:2007:SDM

[JP07b] Willem Jonker and Milan Petković, editors. Secure Data Management: 4th VLDB Workshop, SDM 2007, Vienna, Austria, September 23–24, 2007. Proceedings, volume 4721 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75247-1 (print), 3-540-75248-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75248-6.

Jonker:2008:SDM

[JP08] Willem Jonker and Milan Petković, editors. Secure Data Management: 5th VLDB Workshop, SDM 2008, Auckland, New Zealand, August 24, 2008. Proceedings, volume 5159 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85258-1 (print), 3-540-85259-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
85259-9.

Jiang:2009:CAI

[JP09a] Xiaoyi Jiang and Nicolai Petkov, editors. Computer Analysis of Images and Patterns: 13th International Conference, CAIP 2009, Münster, Germany, September 2– 4, 2009. Proceedings, volume 5702 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03766-6 (print), 3-642-03767-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03767-2.

Jonker:2009:SDM

[JP09b] Willem Jonker and Milan Petković, editors. Secure Data Management: 6th VLDB Workshop, SDM 2009, Lyon, France, August 28, 2009. Proceedings, volume 5776 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04218-X (print), 3-642-04219-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04219-5.

Jonker:2010:SDM

[JP10] Willem Jonker and Milan Petković, editors. Secure Data Management: 7th VLDB Workshop, SDM 2010, Singapore, September 17, 2010. Proceedings, volume 6358 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15545-6 (print), 3-642-15546-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15546-8.

Johnson:2011:AMS

[JP11a] Michael Johnson and Dusko Pavlovic, editors. Algebraic Methodology and Software Technology: 13th International Conference, AMAST 2010, Lac-Beauport, QC, Canada, June 23–25, 2010. Revised Selected Papers, volume 6486 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17795-6 (print), 3-642-17796-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17796-5.

Jonker:2011:SDM

[JP11b] Willem Jonker and Milan Petković, editors. Secure Data Management: 8th VLDB Workshop, SDM 2011, Seattle, WA, USA, September 2, 2011, Proceedings, volume 6933 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23555-7 (print), 3-642-23556-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23556-6.

Jonker:2012:SDM

[JP12a] Willem Jonker and Milan Petković, editors. Secure Data Management: 9th VLDB Workshop, SDM 2012, Istanbul, Turkey, August 27, 2012. Proceedings, volume 7482 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32872-5 (print), 3-642-32873-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32873-2.

Juels:2012:RSP

[JP12b] Ari Juels and Christof Paar, editors. RFID. Security and Privacy: 7th International Workshop, RFIDSec 2011, Amherst, USA, June 26–28, 2011, Revised Selected Papers, volume 7055 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25285-0 (print), 3-642-25286-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25286-0.

Jacob:2005:AIS

[JPBT05] Christian Jacob, Marcin L.

Pilat, Peter J. Bentley, and Jonathan Timmis, editors. Artificial immune systems: 4th international conference, ICARIS 2005, Banff, Alberta, Canada, August 14–17, 2005: proceedings, volume 3627 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28175-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.875 .I23 2005. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3627;   http:
//www.springerlink.
com/content/978-3-540-
28175-7.

Jensen:2004:TMS

[JPD04] Christian Jensen, Stefan Poslad, and Theo Dimitrakos, editors. Trust Management: Second International Conference, iTrust 2004 Oxford, UK, March 29–April 1, 2004 Proceedings, volume 2995 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21312-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2995.htm;
 http://www.springerlink.
com/content/978-3-540-
21312-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2995;
 http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b96545.

Jonoska:2004:AMC

[JPR04] Nataša Jonoska, Gheorghe Păun, and Grzegorz Rozenberg, editors. Aspects of Molecular Computing: Essays Dedicated to Tom Head on the Occasion of His 70th Birthday, volume 2950 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20781-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.887 .A78 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2950.htm;
http://www.springerlink.
com/content/978-3-540-
20781-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2950;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94864.

Jin:2004:GCCa

[JPXS04a] Hai Jin, Yi Pan, Nong Xiao, and Jianhua Sun, editors. Grid and Cooperative Computing— GCC 2004: Third International Conference, Wuhan, China, October 21–24, 2004, Proceedings, volume 3251 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23564-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3251.htm;
 http://www.springerlink.
com/content/978-3-540-
23564-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3251;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100774.

Jin:2004:GCCb

[JPXS04b] Hai Jin, Yi Pan, Nong Xiao, and Jianhua Sun, editors. Grid and Cooperative Computing— GCC 2004 Workshops: GCC 2004 International Workshops, IGKG, SGT Giss, AAC-GEVO, and WS, Wuhan, China, October 21–24, 2004. Proceedings, volume 3252 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23578-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3252.htm;
 http://www.springerlink.
com/content/978-3-540-
23578-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3252;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100775.

Joye:2004:CHE

[JQ04] Marc Joye and Jean-Jacques Quisquater, editors. Cryptographic Hardware and Embedded Systems—CHES 2004: 6th International Workshop Cambridge, MA, USA, August 11–13, 2004. Proceedings, volume 3156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22666-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3156.htm;
 http://www.springerlink.
com/content/978-3-540-
22666-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=3156;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99451.

Jansen:1998:AAC

[JR98] Klaus Jansen and Jose D. P. Rolim, editors. Approximation algorithms for combinatorial optimization: International Workshop, APPROX ’98, Aalborg, Denmark, July 18–19, 1998: Proceedings, volume 1444 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64736-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A65 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1444.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64736-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1444.

Selected papers from AP-PROX ’98, held July 18– 19, 1998, University of Aalborg, Denmark, in conjunction with ICALP ’98.

Jin:2005:NPC

[JRJ05] Hai Jin, Daniel A. Reed, and Wenbin Jiang, editors. Network and parallel computing: IFIP international conference, NPC 2005, Beijing, China, November 30 - December 3, 2005: proceedings, volume 3779 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29810-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .N73 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29810-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3779.

Janowicz:2009:GST

[JRL09] Krzysztof Janowicz, Martin Raubal, and Sergei Levashkin, editors. GeoSpatial Semantics: Third International Conference, GeoS 2009, Mexico City, Mexico, December 3– 4, 2009. Proceedings, volume 5892 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10435-5 (print), 3-642-10436-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10436-7.

Jin:2007:AAP

[JRPP07] Hai Jin, Omer F. Rana, Yi Pan, and Viktor K. Prasanna, editors. Algorithms and Architectures for Parallel Processing: 7th International Conference, ICA3PP 2007, Hangzhou, China, June 11–14, 2007. Proceedings, volume 4494 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72904-6 (print), 3-540-72905-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72905-1.

Junger:2003:COE

[JRR03] M. (Michael) Jünger, G. (Gerhard) Reinelt, and Giovanni Rinaldi, editors. Combinatorial optimization — Eureka, you shrink!: papers dedicated to Jack Edmonds: 5th International Workshop, Aussois, France, March 5– 9, 2001: Revised Papers, volume 2570 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00580-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA402.5 .C54515 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2570.htm;
 http://www.springerlink.
com/content/978-3-540-
00580-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2570.

Also available via the World Wide Web.

Jacobson:2009:SAC

[JRSN09] Michael J. Jacobson Jr., Vincent Rijmen, and Reihaneh Safavi-Naini, editors. Selected Areas in Cryptography: 16th Annual International Workshop, SAC 2009, Calgary, Alberta, Canada, August 13–14, 2009, Revised Selected Papers, volume 5867 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05443-9 (print), 3-642-05445-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        05445-7.

Joseph:1984:FST

[JS84] M. Joseph and Rudrapatna Shyamasundar, editors. Foundations of software technology and theoretical computer science: fourth conference, Bangalore, India, Dec. 13–15, 1984, proceedings, volume 181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-13883-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .F686 1984. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0181.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        13883-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=181.

Joyce:1994:HOL

[JS94] Jeffrey J. Joyce and Carl-Johan H. Seger, editors. Higher order logic theorem proving and its applications: 6th International Workshop, HUG ’93, Vancouver, B.C., Canada, August 11–13, 1993: proceedings, volume 780 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57826-9 (New York), 3-540-57826-9 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 H54 1994. DM96.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0780.
htm;   http://www.springerlink.
com/content/978-0-387-
57826-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=780.

Papers presented at the 1993 HOL User’s Group Workshop.

Jazayeri:1997:SEE

[JS97] Mehdi Jazayeri and Helmut Schauer, editors. Software engineering: ESEC/FSE ’97: 6th European Software Engineering Conference held jointly with the 5th ACM SIGSOFT Symposium on the Foundations of Software Engineering, Zürich, Switzerland, September 22–25, 1997: proceedings, volume 1301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63531-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .S665 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1301.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63531-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1301.

Consists of 27 revised full papers selected from a total of 194 submissions.

Jantke:2001:DSI

[JS01] K. P. Jantke and Ayumi Shinohara, editors. Discovery Science: 4th International Conference, DS 2001, Washington, DC, USA, November 25–28, 2001. Proceedings, volume 2226 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42956-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q174 .D57 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2226.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42956-2;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2226.

Jonoska:2002:DCI

[JS02] Nataša Jonoska and Nadrian C. Seeman, editors. DNA computing: 7th International Workshop on DNA-Based Computers, DNA 7, Tampa, FL, USA, June 10–13, 2001: Revised Papers, volume 2340 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43775-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.887 .I58 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2340.htm;
http://www.springerlink.
com/content/978-3-540-
43775-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2340.

Jedrzejowicz:2005:MFC

[JS05] Joanna Jedrzejowicz and Andrzej Szepietowski, editors. Mathematical foundations of computer science 2005: 30th international symposium, MFCS 2005, Gdansk, Poland, August 29– september 2, 2005, proceedings, volume 3618 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28702-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28702-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3618.

Jedlitschka:2008:PFS

[JS08] Andreas Jedlitschka and Outi Salo, editors. Product-Focused Software Process Improvement: 9th International Conference, PROFES 2008 Monte Porzio Catone, Italy, June 23–25, 2008 Proceedings, volume 5089 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69564-8 (print), 3-540-69566-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69566-0.

Jhala:2011:VMC

[JS11a] Ranjit Jhala and David Schmidt, editors. Verification, Model Checking, and Abstract Interpretation: 12th International Conference, VMCAI 2011, Austin, TX, USA, January 23– 25, 2011. Proceedings, volume 6538 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18274-7 (print), 3-642-18275-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18275-4.

Jouannaud:2011:CPP

[JS11b] Jean-Pierre Jouannaud and Zhong Shao, editors. Certified Programs and Proofs: First International Conference, CPP 2011, Kenting, Taiwan, December 7–9, 2011. Proceedings, volume 7086 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25378-4 (print), 3-642-25379-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25379-9.

JardimNunes:2005:UML

[JSdSA05] Nuno Jardim Nunes, Bran Selic, Alberto Rodrigues da Silva, and Ambrosio Toval Alvarez, editors. UML modeling languages and applications: <<UML>> 2004 satellite activities, Lisbon, Portugal, October 11–15, 2004: revised selected papers, volume 3297 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25081-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 U45 2004b. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25081-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3297.

Jha:2010:RAI

[JSK10] Somesh Jha, Robin Sommer, and Christian Kreibich, editors. Recent Advances in Intrusion Detection: 13th International Symposium, RAID 2010, Ottawa, Ontario, Canada, September 15–17, 2010. Proceedings, volume 6307 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15511-1 (print), 3-642-15512-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15512-3.

Jansen:2004:AOA

[JSO04] Klaus Jansen and Roberto Solis-Oba, editors. Approximation and Online Algorithms: First International Workshop, WAOA 2003, Budapest, Hungary, September 16– 18, 2003: Revised Papers, volume 2909 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21079-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN viii + 268. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2909.htm;
 http://www.springerlink.
com/content/978-3-540-
21079-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2909;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95598.

Jansen:2011:AOA

[JSO11] Klaus Jansen and Roberto Solis-Oba, editors. Approximation and Online Algorithms: 8th International Workshop, WAOA 2010, Liverpool, UK, September 9–10, 2010. Revised Papers, volume 6534 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18317-4 (print), 3-642-18318-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18318-8.

Jamali:2008:MMA

[JSS08] Nadeem Jamali, Paul Scerri, and Toshiharu Sugawara, editors. Massively Multi-Agent Technology: AAMAS Workshops, MMAS 2006, LS-MAS 2006, and CCMMS 2007 Hakodate, Japan, May 9, 2006 Honolulu, HI, USA, May 15, 2007 Selected and Revised Papers, volume 5043 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85448-7 (print), 3-540-85449-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85449-4.

Jain:2011:SOP

[JSV11] Hemant Jain, Atish P.

Sinha, and Padmal Vitharana, editors. Service-Oriented Perspectives in Design Science Research: 6th International Conference, DESRIST 2011, Milwaukee, WI, USA, May 5–6, 2011. Proceedings, volume 6629 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20632-8 (print), 3-642-20633-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20633-7.

Jeffay:2003:QSI

[JSW03] Kevin Jeffay, Ion Stoica, and Klaus Wehrle, editors. Quality of service-IWQoS 2003: 11th International Workshop, Berkeley, California, USA, June 2–4, 2003: Proceedings, volume 2707 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40281-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I97 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2707.htm;
 http://www.springerlink.
com/content/978-3-540-
40281-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2707.

Also available via the World Wide Web.

Janssen:2011:EGI

[JSWhT11] Marijn Janssen, Hans J. Scholl, Maria A. Wimmer, and Yao hua Tan, editors. Electronic Government: 10th IFIP WG 8.5 International Conference, EGOV 2011, Delft, The Netherlands, August 28 — September 2, 2011. Proceedings, volume 6846 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22877-1 (print), 3-642-22878-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22878-0.

Jantke:1995:ALT

[JSZ95] K. P. Jantke, Takeshi Shinohara, and Thomas Zeugmann, editors. Algorithmic learning theory: 6th international workshop, ALT 95, Tokyo, Japan, October 1995: proceedings, volume 997 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60454-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A48 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0997.
htm;   http://www.springerlink.
com/content/978-3-540-
60454-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=997.

Jarecki:2009:PKC

[JT09] Stanisław Jarecki and Gene Tsudik, editors. Public Key Cryptography — PKC 2009: 12th International Conference on Practice and Theory in Public Key Cryptography, Irvine, CA, USA, March 18–20, 2009. Proceedings, volume 5443 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00467-9 (print), 3-642-00468-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00468-1.

Jorge:2005:KDD

[JTB^{+}05] Alípio M. Jorge, Luís Torgo, Pavel B. Brazdil, Rui Camacho, and João Gama, editors. Knowledge Discovery in Databases: PKDD 2005: 9th European Conference on Principles and Practice of Knowledge Discovery in Databases, Porto, Portugal, October 3–7, 2005. Proceedings, volume 3721 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29244-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29244-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3721.

Juels:2004:FCI

[Jue04] Ari Juels, editor. Financial Cryptography: 8th International Conference, FC 2004, Key West, FL, USA, February 9–12, 2004: Revised Papers, volume 3110 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22420-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3110.htm;
 http://www.springerlink.
com/content/978-3-540-
22420-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3110;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98935.

Jul:1998:EOO

[Jul98] Eric Jul, editor. ECOOP ’98 — object-oriented programming: 12th European Conference, Brussels, Belgium, July 20– 24, 1998: Proceedings, volume 1445 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64737-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1445.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64737-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1445.

Jurgens:2002:ISD

[Jür02] Marcus Jürgens. Index Structures for Data Warehouses, volume 1859 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43368-6. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 132 pp. LCCN QA76.9.D26 J87 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1859.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43368-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1859.

Johnson:2006:AMS

[JV06] Michael Johnson and Varmo Vene, editors. Algebraic Methodology and Software Technology: 11th International Conference, AMAST 2006, Kuressaare, Estonia, July 5–8, 2006. Proceedings, volume 4019 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35633-9 (print), 3-540-35636-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35636-3.

Jonsson:2004:RAI

[JVA04] Erland Jonsson, Alfonso Valdes, and Magnus Almgren, editors. Recent Advances in Intrusion Detection: 7th International Symposium, RAID 2004, Sophia Antipolis, France, September 15–17, 2004, Proceedings, volume 3224 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23123-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 R34 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3224.
htm;   http://www.springerlink.
com/content/978-3-540-
23123-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3224;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100714.

Jensen:2009:TPNa

[JvdA09] Kurt Jensen and Wil M. P. van der Aalst, editors. Transactions on Petri Nets and Other Models of Concurrency II: Special Issue on Concurrency in Process-Aware Information Systems, volume 5460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00898-4 (print), 3-642-00899-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00899-3.

Jensen:2008:TPN

[JvdAB08] Kurt Jensen, Wil M. P. van der Aalst, and Jonathan Billington, editors. Transactions on Petri Nets and Other Models of Concurrency I, volume 5100 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89286-9 (print), 3-540-89287-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89287-8.

Jensen:2012:TPNb

[JvdAM^{+}12] Kurt Jensen, Wil M.

van der Aalst, Marco Ajmone Marsan, Giuliana Franceschinis, Jetty Kleijn, and Lars Michael Kristensen, editors. Transactions on Petri Nets and Other Models of Concurrency VI, volume 7400 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35178-6 (print), 3-642-35179-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35179-2.

Jensen:1974:PUM

[JW74] Kathleen Jensen and Niklaus Wirth. PAS-CAL: user manual and report, volume 18 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 0-387-06950-X. ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 170 pp. LCCN QA267.A1 L43 no.18. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0018.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        06950-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=18.

Jensen:1975:PUM

[JW75] Kathleen Jensen and Niklaus Wirth. PAS-CAL, user manual and report, volume 18 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, second edition, 1975. CODEN LNCSD9. ISBN 0-387-07167-9. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 167 pp. LCCN QA76.73.P35 .J461 1974. US$5.90. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0018.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        07167-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=18.

Jajodia:2005:DAS

[JW05] Sushil Jajodia and Duminda Wijesekera, editors. Data and applications security XIX: 19th Annual IFIP WG 11.3 Working Conference on Data and Applications Security, Storrs, CT, USA, August 7–10, 2005: proceedings, volume 3654 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28138-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D314 I34 2005. URL http:

//www.springerlink.
com/content/978-3-540-
28138-2;   http://www.
springerlink.com/link.
asp?id=2d4phthb763b.

Jiao:2006:ANCa

[JWbG^{+}06] Licheng Jiao, Lipo Wang, Xin bo Gao, Jing Liu, and Feng Wu, editors. Advances in Natural Computation: Second International Conference, ICNC 2006, Xi’an, China, September 24–28, 2006. Proceedings, Part I, volume 4221 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45901-4 (print), 3-540-45902-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        45902-6.

Jiao:2006:ANCb

[JWG^{+}06] Licheng Jiao, Lipo Wang, Xinbo Gao, Jing Liu, and Feng Wu, editors. Advances in Natural Computation: Second International Conference, ICNC 2006, Xi’an, China, September 24– 28, 2006. Proceedings, Part II, volume 4222 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45907-3 (print), 3-540-45909-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        45909-5.

Jia:2006:MAH

[JWH06] Xiaohua Jia, Jie Wu, and Yanxiang He, editors. Mobile Ad-hoc and Sensor Networks: First International Conference, MSN 2005, Wuhan, China, December 13–15, 2005. Proceedings, volume 3794 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30856-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30856-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3794.

Jaffar:1996:CPP

[JY96] Joxan Jaffar and Roland H. C. Yap, editors. Concurrency and parallelism, programming, networking, and security: Second Asian Computing Science Conference, ASIAN ’96, Singapore, December 2–5, 1996: proceedings, volume 1179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-62031-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .A85 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1179.htm;
 http://www.springerlink.
com/content/978-3-540-
62031-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1179.

Jung:2012:ISA

[JY12] Souhwan Jung and Moti Yung, editors. Information Security Applications: 12th International Workshop, WISA 2011, Jeju Island, Korea, August 22–24, 2011. Revised Selected Papers, volume 7115 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27889-2 (print), 3-642-27890-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27890-7.

Joux:2014:SAC

[JY14] Antoine Joux and Amr Youssef, editors. Selected areas in cryptography — SAC 2014: 21st International Conference, Montréal, QC, Canada, August 14–15, 2014: revised selected papers, volume 8781 of Lecture notes in computer science: LNCS sublibrary. SL 4, Security and cryptology. Springer-Verlag Inc., New York, NY, USA, 2014. ISBN 3-319-13050-1 (print), 3-319-13051-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25.

Jakobsson:2004:ACN

[JYZ04] Markus Jakobsson, Moti Yung, and Jianying Zhou, editors. Applied Cryptography and Network Security: Second International Conference, ACNS 2004, Yellow Mountain, China, June 8–11, 2004: Proceedings, volume 3089 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22217-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5102.94 .A28 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3089.
htm;   http://www.springerlink.
com/content/978-3-540-
22217-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3089;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98360.

Jeckle:2003:WSI

[JZ03] Mario Jeckle and Liang-Jie Zhang, editors. Web Services—ICWS-Europe 2003: International Conference ICWS-Europe 2003, Erfurt, Germany, September 23–24, 2003: Proceedings, volume 2853 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20125-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2853.htm;
 http://www.springerlink.
com/content/978-3-540-
20125-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2853;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13818.

Jaatun:2009:CCF

[JZR09] Martin Gilje Jaatun, Gansen Zhao, and Chunming Rong, editors. Cloud Computing: First International Conference, CloudCom 2009, Beijing, China, December 1–4, 2009. Proceedings, volume 5931 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10664-1 (print), 3-642-10665-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10665-1.

Kutvonen:2005:DAI

[KA05] Lea Kutvonen and Nancy Alonistioti, editors. Distributed applications and interoperable systems: 5th IFIP WG 6.1 international conference, DAIS 2005, Athens, Greece, June 15–17, 2005, proceedings, volume 3543 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26262-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26262-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3543;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b137217.

Kannan:2012:DEM

[KA12] Rajkumar Kannan and Frederic Andres, editors. Data Engineering and Management: Second International Conference, ICDEM 2010, Tiruchirappalli, India, July 29–31, 2010. Revised Selected Papers, volume 6411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27871-X (print), 3-642-27872-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27872-3.

Kaagstrom:1998:APC

[Kåg98] Bo Kågström, editor. Applied parallel computing: large scale scientific and industrial problem: 4th International Workshop, PARA ’98, Umeå, Sweden, June 1998: Proceedings, volume 1541 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65414-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P353 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1541.htm;
 http://www.springerlink.
com/content/978-3-540-
65414-8;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1541.

Kahn:1979:SCC

[Kah79] G. Kahn, editor. Semantics of concurrent computation: proceedings of the International Symposium, Evian, France, July 2–4, 1979, volume 70 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CODEN LNCSD9. ISBN 0-387-09511-X, 3-540-09511-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I577 1979. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0070.htm;
http://www.springerlink.
com/content/978-0-387-
09511-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=70.

Kahn:1987:FPL

[Kah87] G. Kahn, editor. Functional programming languages and computer architecture, Portland, Oregon, USA, September 14–16, 1987: proceedings, volume 274 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18317-5 (paperback).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.274. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0274.htm;
 http://www.springerlink.
com/content/978-0-387-
18317-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=274.

Proceedings from the 3rd Conference on Functional Programming Languages and Computer Architecture.

Kahng:2003:INN

[Kah03] Hyun-Kook Kahng, editor. Information Networking: networking technologies for enhanced Internet services: International Conference, ICOIN 2003, Cheju Island, Korea, February 12–14, 2003: revised Selected Papers, volume 2662 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40827-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I5715 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2662.htm;
 http://www.springerlink.

com/content/978-3-540-
40827-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2662;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13389.

Kaiser:2011:LGA

[Kai11] Łukasz Kaiser, editor.

Logic and Games on Automatic Structures: Playing with Quantifiers and Decompositions, volume 6810 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22806-2 (print), 3-642-22807-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 110 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22807-0.

Kaliski:1997:ACC

[Kal97] Burton S. Kaliski, editor. Advances in cryptology, CRYPTO ’97: 17th annual international cryptology conference, Santa Barbara, California, USA, August 17–21, 1997: proceedings, volume 1294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63384-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1294.
htm;   http://www.springerlink.
com/content/978-3-540-
63384-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1294.

Kambayashi:1999:ADT

[Kam99] Y. Kambayashi, editor. Advances in database technologies: ER ’98 Workshops on Data Warehousing and Data Mining, Mobile Data Access, and Collaborative Work Support and Spatio-Temporal Data Management, Singapore, November 19–20, 1998: Proceedings, volume 1552 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65690-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 A346 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1552.htm;
 http://www.springerlink.

com/content/978-3-540-
65690-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1552.

Workshops held at the 17th International Conference on Conceptual Modeling (ER 98).

Kapur:1992:ADC

[Kap92] Deepak Kapur, editor.

Automated deduction, CADE-11: 11th International Conference on Automated Deduction, Saratoga Springs, NY, USA, June 15–18, 1992: proceedings, volume 607 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55602-8 (Berlin), 0-387-55602-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I57 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0607.htm;
http://www.springerlink.
com/content/978-3-540-
55602-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=607.

Kapur:2008:CMA

[Kap08] Deepak Kapur, editor.

Computer Mathematics: 8th Asian Symposium, ASCM 2007, Singapore, December 15–17, 2007. Revised and Invited Papers, volume 5081 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87826-2 (print), 3-540-87827-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87827-8.

Kaplan:2010:ATS

[Kap10] Haim Kaplan, editor. Algorithm Theory — SWAT 2010: 12th Scandinavian Symposium and Workshops on Algorithm Theory, Bergen, Norway, June 21–23, 2010. Proceedings, volume 6139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13730-X (print), 3-642-13731-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13731-0.

Karpinski:1977:FCT

[Kar77] Marek Karpinski, editor.

Fundamentals of computation theory: proceedings of the 1977 International FCT-Conference, Poznan-Kornik, Poland, September 19–23, 1977, volume 56 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CODEN LNCSD9. ISBN 0-387-08442-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1,L43 no.56.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0056.
htm;   http://www.springerlink.
com/content/978-0-387-
08442-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=56.

Karpinski:1983:FCT

[Kar83] Marek Karpinski, editor.

Foundations of computation theory: proceedings of the 1983 International FCT-Conference, Borgholm, Sweden, August 21–27, 1983, volume 158 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CODEN LNCSD9. ISBN 0-387-12689-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I57 1983. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0158.
htm;   http://www.springerlink.
com/content/978-0-387-
12689-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=158.

Organized by the Department of Computer and Information Science, Linköping University and Institute of Technology with the cooperation of the Departments of Computer Science and Mathematics, University of Bonn.

Karagiannis:1991:ISA

[Kar91] D. Karagiannis, editor. Information systems and artificial intelligence: integration aspects: first workshop, Ulm, FRG, March 19–21, 1990, proceedings, volume 474 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-53557-8 (Berlin), 0-387-53557-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I532 1991. DM47.00. URL

http://link.springer-
ny.com/link/service/

series/0558/tocs/t0474.
htm;   http://www.springerlink.
com/content/978-3-540-
53557-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=474.

Papers from the 1st Workshop on Information Systems and Artificial Intelligence, held at FAW Ulm March 19–21, 1990.

Karagiannis:1994:DES

[Kar94] D. Karagiannis, editor. Database and expert systems applications: 5th international conference, DEXA ’94, Athens, Greece, September 7–9, 1994: proceedings, volume 856 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58435-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D35843 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0856.
htm;   http://www.springerlink.
com/content/978-0-387-
58435-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=856.

Karsh:2009:EHA

[Kar09] Ben-Tzion Karsh, editor. Ergonomics and Health Aspects of Work with Computers: International Conference, EHAWC 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19– 24, 2009. Proceedings, volume 5624 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02730-X (print), 3-642-02731-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02731-4.

Krumm:2007:UUC

[KASS07] John Krumm, Gregory D. Abowd, Aruna Seneviratne, and Thomas Strang, editors. Ubi-Comp 2007: Ubiquitous Computing: 9th International Conference, UbiComp 2007, Innsbruck, Austria, September 16–19, 2007. Proceedings, volume 4717 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74852-0 (print), 3-540-74853-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74853-3.

Katwijk:1992:AMT

[Kat92] J. Katwijk, editor. Ada:

moving towards 2000: 11th Ada-Europe International Conference, Zandvoort, The Netherlands, June 1–5, 1992: proceedings, volume 603 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55585-4 (Berlin), 0-387-55585-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 A24 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0603.htm;
http://www.springerlink.
com/content/978-3-540-
55585-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=603.

Katoen:1999:FMR

[Kat99] Joost-Pieter Katoen, editor. Formal methods for real-time and probabilistic systems: 5th International AMAST Workshop, ARTS’99, Bamberg, Germany, May 26–28, 1999: Proceedings, volume 1601 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66010-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1601.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1601.
htm;   http://www.springerlink.
com/content/978-3-540-
66010-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1601.

Kaynak:2003:ANN

[Kay03] Okyay Kaynak, editor. Artificial neural networks and neural information processing — ICANN/ ICONIP 2003: joint International Conference ICANN/ICONIP 2003, Istanbul, Turkey, June 26–29, 2003: Proceedings, volume 2714 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40408-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I23 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2714.htm;

http://www.springerlink.
com/content/978-3-540-
40408-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2714.

Krieg-Bruckner:1992:EES

[KB92] B. Krieg-Bruckner, editor. ESOP ’92: 4th European Symposium on Programming, Rennes, France, February 26– 28, 1992: proceedings, volume 582 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55253-7 (Berlin), 0-387-55253-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0582.htm;
http://www.springerlink.
com/content/978-3-540-
55253-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=582.

Kwon:2005:WWG

[KBC05] Yong-Jin Kwon, Alain Bouju, and Christophe Claramunt, editors. Web and wireless geographical information systems: 4th international workshop, W2GIS 2004, Goyang, Korea, November 26–27, 2004. Revised selected papers, volume 3428 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26004-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26004-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3428;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136505.

Kuflik:2009:AUU

[KBC^{+}09] Tsvi Kuflik, Shlomo Berkovsky, Francesca Carmagnola, Dominikus Heckmann, and Antonio Krüger, editors. Advances in Ubiquitous User Modelling: Revised Selected Papers, volume 5830 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05038-7 (print), 3-642-05039-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
05039-8.

Kotasek:2012:MEM

[KBČ^{+}12] Zdeněk Kotásek, Jan Bouda, Ivana Černá, Lukáš Sekanina, Tomáš Vojnar, and David Antoš, editors. Mathematical and Engineering Methods in Computer Science: 7th International Doctoral Workshop, MEMICS 2011, Lednice, Czech Republic, October 14–16, 2011, Revised Selected Papers, volume 7119 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25928-6 (print), 3-642-25929-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25929-6.

Kundig:1987:ESN

[KBD87] A. Kundig, R. E. Buhrer, and J. Dahler, editors. Embedded systems: new approaches to their formal description and design: an advanced course, Zürich, Switzerland, March 5– 7, 1986, volume 284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-18581-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 E631 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0284.htm;
 http://www.springerlink.
com/content/978-0-387-
18581-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=284.

Klusch:2003:IIA

[KBEP03] Matthias Klusch, Sonia Bergamaschi, Pete Edwards, and Paolo Petta, editors. Intelligent information agents: the AgentLink perspective, volume 2586 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00759-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 273 pp. LCCN QA76.76.I58 I582 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2586.
htm;   http://www.springerlink.
com/content/978-3-540-
00759-3;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2586.

Also available via the World Wide Web.

Kosch:2003:EPP

[KBH03] Harald Kosch, László Böszörményi, and Hermann Hellwagner, editors. Euro-Par 2003 parallel processing: 9th International Euro-Par Conference, Klagenfurt, Austria, August 26– 29, 2003: Proceedings, volume 2790 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40788-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5443 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2790.htm;
http://www.springerlink.
com/content/978-3-540-
40788-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2790.

Kolp:2006:AOI

[KBHSW06] Manuel Kolp, Paolo Bresciani, Brian Henderson-Sellers, and Michael Winikoff, editors. Agent-Oriented Information Systems III: 7th International Bi-Conference Workshop, AOIS 2005, Utrecht, Netherlands, July 26, 2005, and Klagenfurt, Austria, October 27, 2005, Revised Selected Papers, volume 3529 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48291-1 (print), 3-540-48292-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48292-5.

Kermarrec:2007:EPP

[KBP07] Anne-Marie Kermarrec, Luc Bougé, and Thierry Priol, editors. Euro-Par 2007 Parallel Processing: 13th International Euro-Par Conference, Rennes ,France , August 28–31, 2007. Proceedings, volume 4641 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74465-7 (print), 3-540-74466-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74466-5.

Kozen:1984:LPW

[KC84] Dexter Kozen and Edmund Clarke, editors. Logics of Programs, workshop, Carnegie Mellon University, Pittsburgh, PA, June 6–8, 1983, volume 164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-12896-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L594 1983. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0164.htm;
http://www.springerlink.
com/content/978-0-387-
12896-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=164.

Proceedings of the Logics of Programs Workshop.

Koopman:2000:IFL

[KC00] Pieter Koopman and Chris Clack, editors. Implementation of functional languages: 11th International Workshop, IFL’99, Lochem, The Netherlands, September 7–10, 1999: Selected Papers, volume 1868 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67864-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .I56 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1868.htm;
 http://www.springerlink.
com/content/978-3-540-
67864-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1868.

Kontio:2002:SQE

[KC02] Jyrki Kontio and Reidar Conradi, editors. Software Quality—ECSQ 2002: 7th International Conference, Helsinki, Finland, June 9–13, 2002. Proceedings, volume 2349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43749-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.Q35 E89 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2349.htm;
 http://www.springerlink.
com/content/978-3-540-
43749-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2349.

Kamel:2005:IAR

[KC05] Mohamed Kamel and Aurélio Campilho, editors. Image analysis and recognition: second international conference, ICIAR 2005, Toronto, Canada, September 28– 30, 2005, proceedings, volume 3656 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29069-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29069-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3656.

Kamel:2007:IAR

[KC07] Mohamed Kamel and Aurélio Campilho, editors. Image Analysis and Recognition: 4th International Conference, ICIAR 2007, Montreal, Canada, August 22–24, 2007. Proceedings, volume 4633 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74258-1 (print), 3-540-74260-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74260-9.

Kamel:2009:IAR

[KC09] Mohamed Kamel and Aurélio Campilho, editors. Image Analysis and Recognition: 6th International Conference, ICIAR 2009, Halifax, Canada, July 6–8, 2009. Proceedings, volume 5627 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02610-9 (print), 3-642-02611-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02611-9.

Kamel:2011:IARa

[KC11a] Mohamed Kamel and Aurélio Campilho, editors. Image Analysis and Recognition: 8th International Conference, ICIAR 2011, Burnaby, BC, Canada, June 22– 24, 2011. Proceedings, Part I, volume 6753 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21592-0 (print), 3-642-21593-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21593-3.

Kamel:2011:IARb

[KC11b] Mohamed Kamel and Aurélio Campilho, editors. Image Analysis and Recognition: 8th International Conference, ICIAR 2011, Burnaby, BC, Canada, June 22– 24, 2011. Proceedings, Part II, volume 6754 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21595-5 (print), 3-642-21596-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21596-4.

Kutylowski:2009:FCT

[KCG09] Mirosław Kutyłowski, Witold Charatonik, and Maciej Gebala, editors. Fundamentals of Computation Theory: 17th International Symposium, FCT 2009, Wrocław, Poland, September 2– 4, 2009. Proceedings, volume 5699 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03408-X (print), 3-642-03409-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03409-1.

Kutylowski:2008:AAW

[KCK08] Mirosław Kutyłowski, Jacek Cichoń, and Przemysław Kubiak, editors. Algorithmic Aspects of Wireless Sensor Networks: Third International Workshop, ALGO-SENSORS 2007, Wroclaw, Poland, July 14, 2007, Revised Selected Papers, volume 4837 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77870-5 (print), 3-540-77871-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77871-4.

Konstan:2011:UMA

[KCMO11] Joseph A. Konstan, Ricardo Conejo, José L.

Marzo, and Nuria Oliver, editors. User Modeling, Adaption and Personalization: 19th International Conference, UMAP 2011, Girona, Spain, July 11–15, 2011. Proceedings, volume 6787 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22361-3 (print), 3-642-22362-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22362-4.

Kurumatani:2004:MMU

[KCO04] Koichi Kurumatani, Shu-Heng Chen, and Azuma Ohuchi, editors. Multiagent for Mass User Support: International Workshop, MAMUS 2003, Acapulco, Mexico, August 10, 2003: Revised and Invited Papers, volume 3012 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21940-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN H61.3 .M36 2003. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t3012.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21940-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3012;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96978.

Kalker:2004:DWS

[KCR04] Ton Kalker, Ingemar J. Cox, and Yong Man Ro, editors. Digital Watermarking: Second International Workshop, IWDW 2003, Seoul, Korea, October 20–22, 2003: Revised Papers, volume 2939 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21061-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I89 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2939.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        21061-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2939;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95658.

Kang:2008:ACI

[KCYL08] Lishan Kang, Zhihua Cai, Xuesong Yan, and Yong Liu, editors. Advances in Computation and Intelligence: Third International Symposium, ISICA 2008 Wuhan, China, December 19–21, 2008 Proceedings, volume 5370 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92136-2 (print), 3-540-92137-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92137-0.

Kleijn:1999:ATP

[KD99] Jetty Kleijn and Susanna Donatelli, editors. Application and theory of Petri nets 1999: 20th International Conference, ICATPN’99 Williamsburg, Virginia, USA, June 21–25, 1999 proceedings, volume 1639 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66132-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A66 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1639.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66132-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1639.

Khendek:2005:TCS

[KD05] Ferhat Khendek and Rachida Dssouli, editors. Testing of communicating systems: 17th IFIP TC 6/WG 6.1 international conference, TESTCOM 2005, Montreal, Canada, May 31– June, 2005, proceedings, volume 3502 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26054-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26054-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3502;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136676.

Kaschek:2011:ECM

[KD11] Roland Kaschek and Lois Delcambre, editors. The Evolution of Conceptual Modeling: From a Historical Perspective towards the Future of Conceptual Modeling, volume 6520 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17504-X (print), 3-642-17505-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17505-3.

Konig:2011:KBIa

[KDH^{+}11a] Andreas König, Andreas Dengel, Knut Hinkelmann, Koichi Kise, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 15th International Conference, KES 2011, Kaiserslautern, Germany, September 12– 14, 2011, Proceedings, Part I, volume 6881 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23850-5 (print), 3-642-23851-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23851-2.

Konig:2011:KBIb

[KDH^{+}11b] Andreas König, Andreas Dengel, Knut Hinkelmann, Koichi Kise, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 15th International Conference, KES 2011, Kaiserslautern, Germany, September 12– 14, 2011, Proceedings, Part II, volume 6882 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23862-9 (print), 3-642-23863-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23863-5.

Konig:2011:KBIc

[KDH^{+}11c] Andreas König, Andreas Dengel, Knut Hinkelmann, Koichi Kise, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 15th International Conference, KES 2011, Kaiserslautern, Germany, September 12– 14, 2011, Proceedings, Part III, volume 6883 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23853-X (print), 3-642-23854-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23854-3.

Konig:2011:KBId

[KDH^{+}11d] Andreas König, Andreas Dengel, Knut Hinkelmann, Koichi Kise, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 15th International Conference, KES 2011, Kaiserslautern, Germany, September 12– 14, 2011, Proceedings, Part IV, volume 6884 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23865-3 (print), 3-642-23866-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23866-6.

Kirby:2001:POS

[KDS01] Graham N. C. Kirby, Alan Dearle, and Dag I. K. Sjøberg, editors. Persistent object systems: design, implementation, and use: 9th International Workshop, POS-9, Lillehammer, Norway, September 6– 8, 2000: Revised Papers, volume 2135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42735-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2135. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2135.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42735-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2135.

Kwak:2010:ISP

[KDWW10] Jin Kwak, Robert H. Deng, Yoojae Won, and Guilin Wang, editors. Information Security, Practice and Experience: 6th International Conference, IS-PEC 2010, Seoul, Korea, May 12–13, 2010. Proceedings, volume 6047 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12826-2 (print), 3-642-12827-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12827-1.

Kaagstrom:2007:APC

[KEDW07] Bo Kågström, Erik Elmroth, Jack Dongarra, and Jerzy Waśniewski, editors. Applied Parallel Computing. State of the Art in Scientific Computing: 8th International Workshop, PARA 2006, Umeå, Sweden, June 18–21, 2006, Revised Selected Papers, volume 4699 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75754-6 (print), 3-540-75755-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75755-9.

Keidar:2009:DCI

[Kei09] Idit Keidar, editor. Distributed Computing: 23rd International Symposium, DISC 2009, Elche, Spain, September 23–25, 2009. Proceedings, volume 5805 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04354-2 (print), 3-642-04355-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04355-0.

Kerren:2007:HCV

[KEM07] Andreas Kerren, Achim Ebert, and Jörg Meyer, editors. Human-Centered Visualization Environments: GI-Dagstuhl Research Seminar, Dagstuhl Castle, Germany, March 5–8, 2006, Revised Lectures, volume 4417 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71948-2 (print), 3-540-71949-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
71949-6.

Keravnou:1997:AIM

[Ker97] E. T. Keravnou, editor.

Artificial intelligence in medicine: 6th Conference on Artificial Intelligence Europe, AIME ’97, Grenoble, France, March 23–26, 1997: proceedings, volume 1211 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62709-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1211. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1211.htm;
http://www.springerlink.
com/content/978-3-540-
62709-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1211.

Kerckhove:2001:SSM

[Ker01] Michael Kerckhove, editor. Scale-space and morphology in computer vision: Third International Conference, Scale-Space 2001, Vancouver, Canada, July 7–8, 2001: Proceedings, volume 2106 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42317-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .I59 2001; QA267.A1 L43 no.2106. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2016.htm;
 http://www.springerlink.
com/content/978-3-540-
42317-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2106.

Keromytis:2012:FCD

[Ker12] Angelos D. Keromytis, editor. Financial Cryptography and Data Security: 16th International Conference, FC 2012, Kralendijk, Bonaire, February 27–March 2, 2012, Revised Selected Papers, volume 7397 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32945-4 (print), 3-642-32946-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F3 2012. URL http:

//www.springerlink.
com/content/978-3-642-
32946-3.

Kruger:1974:FPK

[KF74] Gerhard Kruger and Rudiger Friehmelt, editors. Fachtagung Prozessrechner 1974, Karlsruhe, 10.–11. Juni 1974. (German) [Technical Conference on Process Computers 1974, Karlsruhe, June 10–11, 1974], volume 12 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CO-DEN LNCSD9. ISBN 3-540-06786-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.12.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0012.
htm;   http://www.springerlink.
com/content/978-3-540-
06786-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=12.

Cover title: GFK-GI-GMR Fachtagung Prozessrechner 1974. Conference sponsored by the Gesellschaft für Kernforschung, the Gesellschaft für Informatik and the VDI/VDE-Gesellschaft Mess- und Regelungstechnik. English or German.

Kuhn:2012:CNL

[KF12] Tobias Kuhn and Norbert E. Fuchs, editors. Controlled Natural Language: Third International Workshop, CNL 2012, Zurich, Switzerland, August 29–31, 2012. Proceedings, volume 7427 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32611-0 (print), 3-642-32612-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32612-7.

Kwan:2011:HLI

[KFfKL11] Reggie Kwan, Joseph Fong, Lam for Kwok, and Jeanne Lam, editors. Hybrid Learning: 4th International Conference, ICHL 2011, Hong Kong, China, August 10–12, 2011. Proceedings, volume 6837 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22762-7 (print), 3-642-22763-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
22763-9.

Kortuem:2007:SSC

[KFLS07] Gerd Kortuem, Joe Finney, Rodger Lea, and Vasughi Sundramoorthy, editors. Smart Sensing and Context: Second European Conference, EuroSSC 2007, Kendal, England, October 23–25, 2007. Proceedings, volume 4793 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75695-7 (print), 3-540-75696-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75696-5.

Kovacs:2007:RAT

[KFM07] László Kovács, Norbert Fuhr, and Carlo Meghini, editors. Research and Advanced Technology for Digital Libraries: 11th European Conference, ECDL 2007, Budapest, Hungary, September 16– 21, 2007. Proceedings, volume 4675 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74850-4 (print), 3-540-74851-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74851-9.

Kop:2006:NLP

[KFMM06] Christian Kop, Günther Fliedl, Heinrich C. Mayr, and Elisabeth Métais, editors. Natural Language Processing and Information Systems: 11th International Conference on Applications of Natural Language to Information Systems, NLDB 2006, Klagenfurt, Austria, May 31 — June 2, 2006. Proceedings, volume 3999 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34616-3 (print), 3-540-34617-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34617-3.

Koch:2004:WEI

[KFW04] Nora Koch, Piero Fraternali, and Martin Wirsing, editors. Web Engineering: 4th International Conference, ICWE 2004, Munich, Germany, July 26–30, 2004: Proceedings, volume 3140 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22511-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .I28 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3140.
htm;   http://www.springerlink.
com/content/978-3-540-
22511-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3140;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99180.

Kaiser:1974:OSP

[KG74] Claude Kaiser and E. Gelenbe, editors. Operating systems: proceedings of an international symposium held at Rocquencourt, April 23– 25, 1974, volume 16 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CO-DEN LNCSD9. ISBN 0-387-06849-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.16.

URL http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t0016.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        06849-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=16.

English or French. Organized by the Institut de recherche d’informatique et d’automatique.

Kahng:2004:INN

[KG04] Hyun-Kook Kahng and Shigeki Goto, editors. Information Networking: Networking Technologies for Broadband and Mobile Networks: International Conference ICOIN 2004, Busan, Korea, February 18–20, 2004: Revised Selected Papers, volume 3090 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23034-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I57145 2004; QA76.A1 L43 3090. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3090.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23034-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=

0302-9743&volume=3090;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99472.

Kahl:2012:RAM

[KG12] Wolfram Kahl and Timothy G. Griffin, editors. Relational and Algebraic Methods in Computer Science: 13th International Conference, RAMiCS 2012, Cambridge, UK, September 17–20, 2012. Proceedings, volume 7560 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33313-3 (print), 3-642-33314-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33314-9.

Kloos:2011:TUL

[KGG^{+}11] Carlos Delgado Kloos, Denis Gillet, Raquel M. Crespo García, Fridolin Wild, and Martin Wolpers, editors. Towards Ubiquitous Learning: 6th European Conference of Technology Enhanced Learning, EC-TEL 2011, Palermo, Italy, September 20–23, 2011. Proceedings, volume 6964 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23984-6 (print), 3-642-23985-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23985-4.

Katsikas:2004:PKI

[KGL04] Sokratis K. Katsikas, Stefanos Gritzalis, and Javier Lopez, editors. Public Key Infrastructure: First European PKI Workshop: Research and Applications, EuroPKI 2004, Samos Island, Greece, June 25–26, 2004: Proceedings, volume 3093 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22216-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E973 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3093.
htm;   http://www.springerlink.
com/content/978-3-540-
22216-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3093;

http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98201.

Keller:2010:RAM

[KGRD10] Rainer Keller, Edgar Gabriel, Michael Resch, and Jack Dongarra, editors. Recent Advances in the Message Passing Interface: 17th European MPI Users’ Group Meeting, EuroMPI 2010, Stuttgart, Germany, September 12– 15, 2010. Proceedings, volume 6305 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15645-2 (print), 3-642-15646-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15646-5.

Kounev:2008:PEM

[KGS08] Samuel Kounev, Ian Gorton, and Kai Sachs, editors. Performance Evaluation: Metrics, Models and Benchmarks: SPEC International Performance Evaluation Workshop, SIPEW 2008, Darmstadt, Germany, June 27–28, 2008. Proceedings, volume 5119 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69813-2 (print), 3-540-69814-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69814-2.

Kim:2009:ARF

[KGV^{+}09] Jong-Hwan Kim, Shuzhi Sam Ge, Prahlad Vadakkepat, Norbert Jesse, Abdullah Al Manum, Sadasivan Puthusserypady K, Ulrich Rückert, Joaquin Sitte, Ulf Witkowski, Ryohei Nakatsu, Thomas Braunl, Jacky Baltes, John Anderson, Ching-Chang Wong, Igor Verner, and David Ahlgren, editors. Advances in Robotics: FIRA RoboWorld Congress 2009, Incheon, Korea, August 16–20, 2009. Proceedings, volume 5744 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03982-0 (print), 3-642-03983-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
03983-6.

Koenig:2002:ARA

[KH02] Sven Koenig and Robert Holte, editors. Abstraction, Reformulation and Approximation: 5th International Symposium, SARA 2002, Kananaskis, Alberta, Canada, August 2–4, 2002. Proceedings, volume 2371 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43941-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S27 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2371.htm;
http://www.springerlink.
com/content/978-3-540-
43941-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2371.

Kulkarni:2009:SCL

[KH09] Amba Kulkarni and Gérard Huet, editors. Sanskrit Computational Linguistics: Third International Symposium, Hyderabad, India, January 15–17, 2009. Proceedings, volume 5406 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-93884-2 (print), 3-540-93885-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
93885-9.

Koch:2011:CVAa

[KH11a] Reinhard Koch and Fay Huang, editors. Computer Vision — ACCV 2010 Workshops: ACCV 2010 International Workshops, Queenstown, New Zealand, November 8–9, 2010, Revised Selected Papers, Part I, volume 6468 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22821-6 (print), 3-642-22822-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22822-3.

Koch:2011:CVAb

[KH11b] Reinhard Koch and Fay Huang, editors. Computer Vision — ACCV 2010 Workshops: ACCV 2010 International Workshops, Queenstown, New Zealand, November 8–9, 2010, Revised Selected Papers, Part II, volume 6469 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22818-6 (print), 3-642-22819-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22819-3.

Kraemer:2011:SAM

[KH11c] Frank Alexander Kraemer and Peter Herrmann, editors. System Analysis and Modeling: About Models: 6th International Workshop, SAM 2010, Oslo, Norway, October 4–5, 2010, Revised Selected Papers, volume 6598 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21651-X (print), 3-642-21652-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21652-7.

Kuipers:2012:SOS

[KH12] Fernando A. Kuipers and Poul E. Heegaard, editors. Self-Organizing Systems: 6th IFIP TC 6 International Workshop, IWSOS 2012, Delft, The Netherlands, March 15– 16, 2012. Proceedings, volume 7166 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28582-1 (print), 3-642-28583-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28583-7.

Klette:2001:MIA

[KHG01] Reinhard Klette, Thomas S. Huang, and Georgii L’vovich Gimel’farb, editors. Multiimage analysis: 10th International Workshop on Theoretical Foundations of Computer Vision, Dagstuhl Castle, Germany, March 12– 17, 2000: Revised Papers, volume 2032 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42122-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .I557 2000; QA267.A1 L43 no.2032. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2032.htm;
http://www.springerlink.
com/content/978-3-540-
42122-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2032.

Koucheryavy:2006:NGT

[KHI06] Yevgeni Koucheryavy, Jarmo Harju, and Villy B. Iversen, editors. Next Generation Teletraffic and Wired/Wireless Advanced Networking: 6th International Conference, NEW2AN 2006, St. Petersburg, Russia, May 29 — June 2, 2006. Proceedings, volume 4003 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34429-2 (print), 3-540-34430-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34430-8.

Khosla:2005:KBIb

[KHJ05a] Rajiv Khosla, Robert J.

Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems, volume 3682 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28895-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28895-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3682.

Khosla:2005:KBIc

[KHJ05b] Rajiv Khosla, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems, volume 3683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28896-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28896-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3683.

Khosla:2005:KBId

[KHJ05c] Rajiv Khosla, Robert J.

Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems, volume 3684 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28897-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28897-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3684.

Khosla:2005:KBIa

[KHJ05d] Rajiv Khosla, Robert J.

Howlett, and Lakhmi C. Jain, editors. Knowledgebased intelligent information and engineering systems: 9th international conference, KES 2005, Melbourne, Australia, September 14– 16, 2005, proceedings, Part I, volume 3681 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28894-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7801 .I5155 2005; TK7801 .I493 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28894-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3681.

Kowalczyk:2008:SOC

[KHK^{+}08] Ryszard Kowalczyk, Michael Huhns, Matthias Klusch, Zakaria Maamar, and Quoc Bao Vo, editors. Service-Oriented Computing: Agents, Semantics, and Engineering: AAMAS 2008 International Workshop, SO-CASE 2008, Estoril, Portugal, May 12, 2008 Proceedings, volume 5006 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79967-2 (print), 3-540-79968-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79968-9.

Kolfschoten:2010:CTI

[KHL10] Gwendolyn Kolfschoten, Thomas Herrmann, and Stephan Lukosch, editors. Collaboration and Technology: 16th International Conference, CRIWG 2010, Maastricht, The Netherlands, September 20–23, 2010. Proceedings, volume 6257 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15713-0 (print), 3-642-15714-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15714-1.

Klusch:2007:CIA

[KHPS07] Matthias Klusch, Koen V.

Hindriks, Mike P. Papazoglou, and Leon Sterling, editors. Cooperative Information Agents XI: 11th International Workshop, CIA 2007, Delft, The Netherlands, September 19–21, 2007. Proceedings, volume 4676 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75118-1 (print), 3-540-75119-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75119-9.

Koucheryavy:2007:NGT

[KHS07] Yevgeni Koucheryavy, Jarmo Harju, and Alexander Sayenko, editors. Next Generation Teletraffic and Wired/Wireless Advanced Networking: 7th International Conference, NEW2AN 2007 St. Petersburg, Russia, September 10–14, 2007 Proceedings, volume 4712 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74832-6 (print), 3-540-74833-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74833-5.

Kolp:2008:AOI

[KHSM^{+}08] Manuel Kolp, Brian Henderson-Sellers, Haralambos Mouratidis, Alessandro Garcia, Aditya K. Ghose, and Paolo Bresciani, editors. Agent-Oriented Information Systems IV: 8th International Bi-Conference Workshop, AOIS 2006, Hakodate, Japan, May 9, 2006 and Luxembourg, Luxembourg, June 6, 2006, Revised Selected Papers, volume 4898 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77989-2 (print), 3-540-77990-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77990-2.

Konig:2003:FTN

[KHW03] Hartmut König, Monika Heiner, and Adam Wolisz, editors. Formal Techniques for Networked and Distributed Systems— FORTE 2003: 23rd IFIP WG 6.1 International Conference, Berlin, Germany, September 29– October 2, 2003: Proceedings, volume 2767 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20175-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 F67 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2767.
htm;   http://www.springerlink.
com/content/978-3-540-
20175-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2767;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
11965.

Kastens:1982:GPC

[KHZ82] Uwe Kastens, Brigitte Hutt, and Erich Zimmermann. GAG, a practical compiler generator, volume 141 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 3-540-11591-9 (paperback: U.S). ISSN 0302-9743 (print), 1611-3349 (electronic). iv + 156 pp. LCCN QA76.6 .K347. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0141.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        11591-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=141.

Kern-Isberner:2001:CNR

[KI01] Gabriele Kern-Isberner. Conditionals in Nonmonotonic Reasoning and Belief Revision: Considering Conditionals as Agents, volume 2087 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42367-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 190 pp. LCCN QA267.A1 L43 no.2087; Q339.2 .K47 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2087.htm;
http://www.springerlink.
com/content/978-3-540-
42367-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2087.

Kosseim:2012:AAI

[KI12] Leila Kosseim and Diana Inkpen, editors. Advances in Artificial Intelligence: 25th Canadian Conference on Artificial Intelligence, Canadian AI 2012, Toronto, ON, Canada, May 28– 30, 2012. Proceedings, volume 7310 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30352-8 (print), 3-642-30353-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30353-1.

Kiayias:2011:TCC

[Kia11] Aggelos Kiayias, editor. Topics in Cryptology

— CT-RSA 2011: The Cryptographers’ Track at the RSA Conference 2011, San Francisco, CA, USA, February 14–18, 2011. Proceedings, volume 6558 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19073-1 (print), 3-642-19074-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19074-2.

Kienzle:2012:MSE

[Kie12] Jörg Kienzle, editor. Models in Software Engineering: Workshops and Symposia at MOD-ELS 2011, Wellington, New Zealand, October 16–21, 2011, Reports and Revised Selected Papers, volume 7167 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29644-0 (print), 3-642-29645-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29645-1.

Kilian:2001:ACC

[Kil01] Joe Kilian, editor. Advances in cryptology — CRYPTO 2001: 21st Annual International Cryptology Conference, Santa Barbara, California, USA, August 19–23, 2001: Proceedings, volume 2139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42456-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 2001; QA267.A1 L43 no.2139. UK£47.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2139.htm;
http://www.springerlink.
com/content/978-3-540-
42456-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2139.

Kilian:2005:TCS

[Kil05] Joe Kilian, editor. Theory of cryptography: Second theory of cryptography conference, TCC 2005, Cambridge, MA, USA, February 10–12, 2005, proceedings, volume 3378 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24573-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 T44 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24573-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3378;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106171.

Kitagawa:2010:DSAa

[KILW10a] Hiroyuki Kitagawa, Yoshiharu Ishikawa, Qing Li, and Chiemi Watanabe, editors. Database Systems for Advanced Applications: 15th International Conference, DAS-FAA 2010, Tsukuba, Japan, April 1–4, 2010, Proceedings, Part I, volume 5981 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12025-3 (print), 3-642-12026-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12026-8.

Kitagawa:2010:DSAb

[KILW10b] Hiroyuki Kitagawa, Yoshiharu Ishikawa, Qing Li, and Chiemi Watanabe, editors. Database Systems for Advanced Applications: 15th International Conference, DAS-FAA 2010, Tsukuba, Japan, April 1–4, 2010, Proceedings, Part II, volume 5982 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12097-0 (print), 3-642-12098-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12098-5.

Kim:2001:PKC

[Kim01a] Kwangjo Kim, editor.

Public key cryptography: 4th International Workshop on Practice and Theory in Public Key Cryptosystems, PKC 2001, Cheju Island, Korea, February 13–15, 2001: Proceedings, volume 1992 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41658-7 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I567 2001; QA267.A1 L43 no.1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1992.htm;
 http://www.springerlink.
com/content/978-3-540-
41658-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1992.

Kim:2001:HSI

[Kim01b] Won Kim, editor. The Human Society and the Internet: Internet-related socio-economic issues, First International Conference Human.Society@Internet 2001, Seoul, Korea, July 4–6, 2001: Proceedings, volume 2105 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42313-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 I5345 2001; QA267.A1 L43 no.2105. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2105.htm;
 http://www.springerlink.
com/content/978-3-540-
42313-3;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2105.

Kim:2002:ISC

[Kim02] Kwangjo Kim, editor. Information security and cryptology — ICISC 2001: 4th International Conference, Seoul, Korea, December 6–7, 2001: Proceedings, volume 2288 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43319-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I32 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2288.htm;
http://www.springerlink.
com/content/978-3-540-
43319-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2288.

Kim:2005:INC

[Kim05a] Cheeha Kim, editor. Information networking: convergence in broadband and mobile networking. international conference, ICOIN 2005, Jeju Island, Korea, January 31– February 2, 2005. Proceedings, volume 3391 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24467-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I5715 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24467-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3391;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105584.

Kim:2005:AIS

[Kim05b] Tag Gon Kim, editor. Artificial Intelligence and Simulation: 13th International Conference on AI, Simulation, Planning in High Autonomy Systems, AIS 2004, Jeju Island, Korea, October 4–6, 2004. Revised selected papers, volume 3397 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24476-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65. URL

http://www.springerlink.
com/content/978-3-540-
24476-9;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3397;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105609.

Kim:2012:ISC

[Kim12] Howon Kim, editor. Information Security and Cryptology — ICISC 2011: 14th International Conference, Seoul, Korea, November 30 — December 2, 2011. Revised Selected Papers, volume 7259 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31911-4 (print), 3-642-31912-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31912-9.

King:2008:LBP

[Kin08] Andy King, editor.

Logic-Based Program Synthesis and Transformation: 17th International Symposium, LOP-STR 2007, Kongens Lyngby, Denmark, August 23–24, 2007, Revised Selected Papers, volume 4915 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78768-2 (print), 3-540-78769-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78769-3.

Kirchner:1993:RTA

[Kir93] Claude Kirchner, editor. Rewriting techniques and applications: 5th International Conference, RTA-93, Montréal, Canada, June 16–18, 1993: proceedings, volume 690 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56868-9 (Berlin), 0-387-56868-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I555 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0690.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        56868-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=690.

Kirchner:1996:TAP

[Kir96] H. (Helene) Kirchner, editor. Trees in algebra and programming— CAAP ’96: 21st international colloquium, Linköping, Sweden, April 22–24, 1996, proceedings, volume 1059 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61064-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166.2 .C3 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1059.htm;
 http://www.springerlink.
com/content/978-3-540-
61064-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1059.

Kern-Isberner:2005:CII

[KIRK05] Gabriele Kern-Isberner, Wilhelm Rödder, and Friedhelm Kulmann, editors. Conditionals, information, and inference: international workshop, WCII 2002, Hagen, Germany, May 13– 15, 2002, revised selected papers, volume 3301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25332-7 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25332-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3301;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b107184.

Kistler:1995:DOD

[Kis95] James Jay Kistler. Disconnected operation in a distributed file system, volume 1002 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60627-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 249 pp. LCCN QA76.9.D5K57 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1002.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60627-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1002.

Kittler:1988:PRI

[Kit88] Josef Kittler, editor. Pattern recognition: 4th international conference, Cambridge, UK, March 28–30, 1988: proceedings, volume 301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19036-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q327 .P3651 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0301.htm;
 http://www.springerlink.
com/content/978-0-387-
19036-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=301.

“Papers presented at the BPRA 4th International Conference on Pattern Recognition”–Pref. Spine title: Pattern recognition 1988.

Kitano:1998:RRS

[Kit98] Hiroaki Kitano. RoboCup-97: robot soccer World Cup I, volume 1395 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64473-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). various pp. LCCN TJ210.3.R615 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1395.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64473-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1395.

Kaplan:1988:CTR

[KJ88] S. Kaplan and Jean-Pierre Jouannaud, editors. Conditional term rewriting systems: 1st international workshop, Orsay, France, July 8–10, 1987: proceedings, volume 308 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19242-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .C59331 1988. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0308.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        19242-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=308.

Kelemen:1991:FAI

[KJ91] Jozef Kelemen and Ph. Jorrand, editors. Fundamentals of artificial intelligence research: International Workshop FAIR ’91, Smolenice, Czechoslovakia, September 3–13, 1991: proceedings, volume 535 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54507-7, 0-387-54507-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I58 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0535.htm;
 http://www.springerlink.
com/content/978-3-540-
54507-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=535.

Karagiannis:2009:KSE

[KJ09] Dimitris Karagiannis and Zhi Jin, editors. Knowledge Science, Engineering and Management: Third International Conference, KSEM 2009, Vienna, Austria, November 25–27, 2009. Proceedings, volume 5914 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10487-8 (print), 3-642-10488-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10488-6.

Kreibich:2010:DIM

[KJ10] Christian Kreibich and Marko Jahnke, editors. Detection of Intrusions and Malware, and Vulnerability Assessment: 7th International Conference, DIMVA 2010, Bonn, Germany, July 8–9, 2010. Proceedings, volume 6201 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14214-1 (print), 3-642-14215-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14215-4.

Kirda:2009:RAI

[KJB09] Engin Kirda, Somesh Jha, and Davide Balzarotti, editors. Recent Advances in Intrusion Detection: 12th International Symposium, RAID 2009, Saint-Malo, France, September 23–25, 2009. Proceedings, volume 5758 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04341-0 (print), 3-642-04342-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04342-0.

Kinny:2011:APA

[KjHGG11] David Kinny, Jane Yung jen Hsu, Guido Governatori, and Aditya K. Ghose, editors. Agents in Principle, Agents in Practice: 14th International Conference, PRIMA 2011, Wollongong, Australia, November 16–18, 2011. Proceedings, volume 7047 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25043-2 (print), 3-642-25044-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25044-6.

Kanade:2005:AVB

[KJR05] Takeo Kanade, Anil Jain, and Nalini K. Ratha, editors. Audio- and Video-Based Biometric Person Authentication: 5th International Conference, AVBPA 2005, Hilton Rye Town, NY, USA, July 20–22, 2005. Proceedings, volume 3546 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27887-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.S65 I565 2005.

URL http://www.springerlink.
com/content/978-3-540-
27887-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3546.

Kunii:2001:CMI

[KJS01] Hideko S. Kunii, Sushil Jajodia, and Arne Sølvberg, editors. Conceptual Modeling—ER 2001: 20th International Conference on Conceptual Modeling, Yokohama, Japan, November 27–30, 2001. Proceedings, volume 2224 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42866-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26 I567 2001. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t2224.htm;
http://www.springerlink.
com/content/978-3-540-
42866-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2224.

Kelemenova:1987:TTP

[KK87] Alica Kelemenova and Jozef Kelemen, editors. Trends, techniques, and problems in theoretical computer science: selected contributions, volume 281 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18535-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I6281 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0281.htm;
http://www.springerlink.
com/content/978-0-387-
18535-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=281.

Kumar:1995:TPC

[KK95] Ramayya Kumar and Thomas Kropf, editors. Theorem provers in circuit design: theory, practice, and experience: second international conference, TPCD ’94, Bad Herrenalb, Germany, September 26–28, 1994: proceedings, volume 901 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59047-1 (Berlin), 0-387-59047-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.65 .T54 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0901.htm;
 http://www.springerlink.
com/content/978-3-540-
59047-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=901.

Kandzia:1997:CIA

[KK97] Peter Kandzia and Matthias Klusch, editors. Cooperative information agents: First International Workshop, CIA ’97, Kiel, Germany, February 26–28, 1997: proceedings, volume 1202 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62591-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q337 .C66 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1202.
htm;   http://www.springerlink.
com/content/978-3-540-
62591-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1202.

Kirchner:1998:ADC

[KK98] Claude Kirchner and H. (Hélène) Kirchner, editors. Automated deduction, CADE-15: 15th International Conference on Automated Deduction, Lindau, Germany, July 5–10, 1998: Proceedings, volume 1421 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64675-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1421. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1421.htm;
http://www.springerlink.
com/content/978-3-540-
64675-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1421.

Klusch:2000:CIA

[KK00] Matthias Klusch and Larry Kerschberg, editors. Cooperative information agents IV: the future of information agents in cyberspace: 4th International Workshop, CIA 2000, Boston, MA, USA, July 7–9, 2000: Proceedings, volume 1860 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67703-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1860. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1860.htm;
 http://www.springerlink.
com/content/978-3-540-
67703-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1860.

Kucera:2007:MFC

[KK07] Luděk Kučera and Antonín Kučera, editors. Mathematical Foundations of Computer Science 2007: 32nd International Symposium, MFCS 2007 Český Krumlov, Czech Republic, August 26–31, 2007 Proceedings, volume 4708 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74455-X (print), 3-540-74456-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74456-6.

Kolb:2009:DID

[KK09a] Andreas Kolb and Reinhard Koch, editors. Dynamic 3D Imaging: DAGM 2009 Workshop, Dyn3D 2009, Jena, Germany, September 9, 2009. Proceedings, volume 5742 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03777-1 (print), 3-642-03778-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03778-8.

Kordon:2009:RST

[KK09b] Fabrice Kordon and Yvon Kermarrec, editors. Reliable Software Technologies — Ada-Europe 2009: 14th Ada-Europe International Conference, Brest, France, June 8–12, 2009. Proceedings, volume 5570 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01923-4 (print), 3-642-01924-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01924-1.

Katoen:2011:CCT

[KK11a] Joost-Pieter Katoen and Barbara König, editors. CONCUR 2011 — Concurrency Theory: 22nd International Conference, CONCUR 2011, Aachen, Germany, September 6–9, 2011. Proceedings, volume 6901 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23216-7 (print), 3-642-23217-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23217-6.

Katoh:2011:WAC

[KK11b] Naoki Katoh and Amit Kumar, editors. WAL-COM: Algorithms and Computation: 5th International Workshop, WALCOM 2011, New Delhi, India, February 18–20, 2011. Proceedings, volume 6552 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19093-6 (print), 3-642-19094-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19094-0.

Kelemen:2011:CCL

[KK11c] Jozef Kelemen and Alica Kelemenová, editors. Computation, Cooperation, and Life: Essays Dedicated to Gheorghe Păun on the Occasion of His 60th Birthday, volume 6610 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19999-2 (print), 3-642-20000-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20000-7.

Kolman:2011:GTC

[KK11d] Petr Kolman and Jan Kratochvíl, editors. Graph-Theoretic Concepts in Computer Science: 37th International Workshop, WG 2011, Teplá Monastery, Czech Republic, June 21– 24, 2011. Revised Papers, volume 6986 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25869-7 (print), 3-642-25870-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25870-1.

Kon:2011:MAI

[KK11e] Fabio Kon and Anne-Marie Kermarrec, editors. Middleware 2011: ACM/IFIP/USENIX 12th International Middleware Conference, Lisbon, Portugal, December 12–16, 2011. Proceedings, volume 7049 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25820-4 (print), 3-642-25821-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
25821-3.

Kudenko:2005:AAM

[KKA05] Daniel Kudenko, Dimitar Kazakov, and Eduardo Alonso, editors. Adaptive agents and multiagent systems II: adaptation and multi-agent learning, volume 3394 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25260-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 A312 2005. URL http:

//www.springerlink.
com/content/978-3-540-
25260-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3394;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106974.

Koppen:2009:ANIa

[KKC09a] Mario Köppen, Nikola Kasabov, and George Coghill, editors. Advances in Neuro-Information Processing: 15th International Conference, ICONIP 2008, Auckland, New Zealand, November 25–28, 2008, Revised Selected Papers, Part I, volume 5506 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02489-0 (print), 3-642-02490-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02490-0.

Koppen:2009:ANIb

[KKC09b] Mario Köppen, Nikola Kasabov, and George Coghill, editors. Advances in Neuro-Information Processing: 15th International Conference, ICONIP 2008, Auckland, New Zealand, November 25–28, 2008, Revised Selected Papers, Part II, volume 5507 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03039-4 (print), 3-642-03040-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03040-6.

Kranzlmuller:2004:RAP

[KKD04] Dieter Kranzlmüller, Péter Kacsuk, and Jack J. Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 11th European PVM/MPI Users’ Group Meeting, Budapest, Hungary, September 19– 22, 2004: proceedings, volume 3241 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23163-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23163-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3241;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100820.

Kok:2007:KDD

[KKdM^{+}07a] Joost N. Kok, Jacek Koronacki, Ramon Lopez de Mantaras, Stan Matwin, Dunja Mladenič, and Andrzej Skowron, editors. Knowledge Discovery in Databases: PKDD 2007: 11th European Conference on Principles and Practice of Knowledge Discovery in Databases, Warsaw, Poland, September 17–21, 2007. Proceedings, volume 4702 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74975-6 (print), 3-540-74976-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74976-9.

Kok:2007:MLE

[KKdM^{+}07b] Joost N. Kok, Jacek Koronacki, Raomon Lopez de Mantaras, Stan Matwin, Dunja Mladenič, and Andrzej Skowron, editors. Machine Learning: ECML 2007: 18th European Conference on Machine Learning, Warsaw, Poland, September 17–21, 2007. Proceedings, volume 4701 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74957-8 (print), 3-540-74958-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74958-5.

Kranzlmuller:2002:RAP

[KKDV02] Dieter Kranzlmüller, Peter Kacsuk, Jack Dongarra, and Jens Volkert, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 9th European PVM/MPI Users’ Group Meeting, Linz, Austria, September 29–October 2, 2002. Proceedings, volume 2474 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44296-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E975 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2474.htm;
http://www.springerlink.
com/content/978-3-540-
44296-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2474.

Also available via the World Wide Web.

Kamel:2011:AIS

[KKGK11] Mohamed Kamel, Fakhri Karray, Wail Gueaieb, and Alaa Khamis, editors. Autonomous and Intelligent Systems: Second International Conference, AIS 2011, Burnaby, BC, Canada, June 22–24, 2011. Proceedings, volume 6752 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21537-8 (print), 3-642-21538-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21538-4.

Kumar:2003:CSAb

[KKGT03a] Vipin Kumar, Vipin Kumar, Marina L. Gavrilova, and Chih Jeng Kenneth Tan, editors. Computational science and its applications: ICCSA 2003, International Conference, Montréal, Canada, May 18–21, 2003: Proceedings, volume 2668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40161-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I12 2003 pt.1-3 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2668.htm;
 http://www.springerlink.
com/content/978-3-540-
40161-2;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2668.

Kumar:2003:CSAc

[KKGT03b] Vipin Kumar, Vipin Kumar, Marina L. Gavrilova, and Chih Jeng Kenneth Tan, editors. Computational science and its applications: ICCSA 2003, International Conference, Montréal, Canada, May 18–21, 2003: Proceedings, volume 2669 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40156-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I12 2003 pt.1-3 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2669.htm;
http://www.springerlink.
com/content/978-3-540-
40156-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2669.

Kumar:2003:CSAa

[KKGT03c] Vipin Kumar, Vipin Kumar, Marina L. Gavrilova, and Chih Jeng Kenneth Tan, editors. Computational science and its applications: ICCSA 2003, International Conference, Montréal, Canada, May 18–21, 2003: Proceedings. Part 1, volume 2667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40155-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I12 2003 pt.1-3 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2667.htm;
 http://www.springerlink.
com/content/978-3-540-
40155-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2667.

Kropatsch:2007:CAI

[KKH07] Walter G. Kropatsch, Martin Kampel, and Allan Hanbury, editors. Computer Analysis of Images and Patterns: 12th International Conference, CAIP 2007, Vienna, Austria, August 27–29, 2007. Proceedings, volume 4673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74271-9 (print), 3-540-74272-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74272-2.

Kim:2009:DWI

[KKH09] Hyoung-Joong Kim, Stefan Katzenbeisser, and Anthony T. S. Ho, editors. Digital Watermarking: 7th International Workshop, IWDW 2008, Busan, Korea, November 10–12, 2008. Selected Papers, volume 5450 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04437-9 (print), 3-642-04438-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04438-0.

Kishino:2005:ECI

[KKKN05] Fumio Kishino, Yoshifumi Kitamura, Hirokazu Kato, and Noriko Nagata, editors. Entertainment computing — ICEC 2005: 4th international conference, Sanda, Japan, September 19–21, 2005: proceedings, volume 3711 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29034-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.E57 I58 2005; QA769.E57 I58 2005.

URL http://www.springerlink.
com/content/978-3-540-
29034-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3711.

Kanazawa:2011:MLB

[KKKS11] Makoto Kanazawa, András Kornai, Marcus Kracht, and Hiroyuki Seki, editors. The Mathematics of Language: 12th Biennial Conference, MOL 12, Nara, Japan, September 6–8, 2011. Proceedings, volume 6878 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23210-8 (print), 3-642-23211-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23211-4.

Kranakis:2012:FAI

[KKL12] Evangelos Kranakis, Danny Krizanc, and Flaminia Luccio, editors. Fun with Algorithms: 6th International Conference, FUN 2012, Venice, Italy, June 4–6, 2012. Proceedings, volume 7288 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30346-3 (print), 3-642-30347-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30347-0.

Karmouch:2004:MAT

[KKM04] Ahmed Karmouch, Larry Korba, and Edmundo R. M. Madeira, editors. Mobility aware technologies and applications: first international workshop, MATA 2004, Florianopolis, Brazil, October 20–22, 2004, proceedings, volume 3284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23423-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
23423-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3284;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b101423.

Koch:2011:RCA

[KKM^{+}11] Andreas Koch, Ram Krishnamurthy, John McAllister, Roger Woods, and Tarek El-Ghazawi, editors. Reconfigurable Computing: Architectures, Tools and Applications: 7th International Symposium, ARC 2011, Belfast, UK, March 23– 25, 2011. Proceedings, volume 6578 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19474-5 (print), 3-642-19475-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19475-7.

Kotagiri:2007:ADC

[KKMN07] Ramamohanarao Kotagiri, P. Radha Krishna, Mukesh Mohania, and Ekawit Nantajeewarawat, editors. Advances in Databases: Concepts, Systems and Applications: 12th International Conference on Database Systems for Advanced Applications, DASFAA 2007, Bangkok, Thailand, April 9–12, 2007. Proceedings, volume 4443 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71702-1 (print), 3-540-71703-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71703-4.

Kauers:2007:TMM

[KKMW07] Manuel Kauers, Manfred Kerber, Robert Miner, and Wolfgang Windsteiger, editors. Towards Mechanized Mathematical Assistants: 14th Symposium, Calculemus 2007, 6th International Conference, MKM 2007, Hagenberg, Austria, June 27–30, 2007. Proceedings, volume 4573 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73083-4 (print), 3-540-73086-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73086-6.

Kaliski:2002:CHE

[KKP02] Burton S. Kaliski Jr., Çetin K. Koç, and Christof Paar, editors. Cryptographic Hardware and Embedded Systems— CHES 2002: 4th International Workshop Redwood Shores, CA, USA, August 13–15, 20. Revised Papers, volume 2523 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00409-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 C454 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2523.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00409-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2523.

Also available via the

World Wide Web.

Kontogiannis:2010:AGT

[KKS10] Spyros Kontogiannis, Elias Koutsoupias, and Paul G. Spirakis, editors. Algorithmic Game Theory: Third International Symposium, SAGT 2010, Athens, Greece, October 18–20, 2010. Proceedings, volume 6386 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16169-3 (print), 3-642-16170-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16170-4.

Kampis:2011:AALa

[KKS11a] George Kampis, István Karsai, and Eörs Szathmáry, editors. Advances in Artificial Life. Darwin Meets von Neumann: 10th European Conference, ECAL 2009, Budapest, Hungary, September 13–16, 2009, Revised Selected Papers, Part I, volume 5777 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21282-4 (print), 3-642-21283-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21283-3.

Kampis:2011:AALb

[KKS11b] George Kampis, István Karsai, and Eörs Szathmáry, editors. Advances in Artificial Life. Darwin Meets von Neumann: 10th European Conference, ECAL 2009, Budapest, Hungary, September 13–16, 2009, Revised Selected Papers, Part II, volume 5778 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21313-8 (print), 3-642-21314-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21314-4.

Kimmel:2011:CVAa

[KKS11c] Ron Kimmel, Reinhard Klette, and Akihiro Sugimoto, editors. Computer Vision — ACCV 2010: 10th Asian Conference on Computer Vision, Queenstown, New Zealand, November 8– 12, 2010, Revised Selected Papers, Part I, volume 6492 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19314-5 (print), 3-642-19315-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19315-6.

Kimmel:2011:CVAb

[KKS11d] Ron Kimmel, Reinhard Klette, and Akihiro Sugimoto, editors. Computer Vision — ACCV 2010: 10th Asian Conference on Computer Vision, Queenstown, New Zealand, November 8– 12, 2010, Revised Selected Papers, Part II, volume 6493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19308-0 (print), 3-642-19309-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19309-5.

Kimmel:2011:CVAc

[KKS11e] Ron Kimmel, Reinhard Klette, and Akihiro Sugimoto, editors. Computer Vision — ACCV 2010: 10th Asian Conference on Computer Vision, Queenstown, New Zealand, November 8– 12, 2010, Revised Selected Papers, Part III, volume 6494 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19317-X (print), 3-642-19318-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19318-7.

Kimmel:2011:CVAd

[KKS11f] Ron Kimmel, Reinhard Klette, and Akihiro Sugimoto, editors. Computer Vision — ACCV 2010: 10th Asian Conference on Computer Vision, Queenstown, New Zealand, November 8– 12, 2010, Revised Selected Papers, Part IV, volume 6495 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19281-5 (print), 3-642-19282-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19282-1.

Keller:2011:FMC

[KKW11] Rainer Keller, David Kramer, and Jan-Philipp Weiss, editors. Facing the Multicore-Challenge: Aspects of New Paradigms and Technologies in Parallel Computing, volume 6310 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16232-0 (print), 3-642-16233-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16233-6.

Keller:2012:FMC

[KKW12] Rainer Keller, David Kramer, and Jan-Philipp Weiss, editors. Facing the Multicore — Challenge II: Aspects of New Paradigms and Technologies in Parallel Computing, volume 7174 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30396-X (print), 3-642-30397-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30397-5.

Karlsson:1988:SSW

[KL88] R. Karlsson and Andrzej Lingas, editors. SWAT 88: 1st Scandinavian Workshop on Algorithm Theory, Halmstad, Sweden, July 5–8, 1988: proceedings, volume 318 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19487-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .S931 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0318.htm;
 http://www.springerlink.
com/content/978-0-387-
19487-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=318.

Kirchner:1992:ALP

[KL92] H. Kirchner and G. Levi, editors. Algebraic and logic programming: third international conference, Volterra, Italy, September 2–4, 1992: proceedings, volume 632 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55873-X (Berlin), 0-387-55873-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .A45842 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0632.htm;
http://www.springerlink.
com/content/978-3-540-
55873-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=632.

Kanchanasut:1995:ACK

[KL95] Kanachana Kanchanasut and Jean-Jacques Levy, editors. Algorithms, concurrency, and knowledge: 1995 Asian Computing Science Conference, ACSC ’95, Pathumthani, Thailand, December 11– 13, 1995: proceedings, volume 1023 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60688-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43A85 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1023.htm;
http://www.springerlink.
com/content/978-3-540-
60688-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1023.

Karlsson:1996:ATS

[KL96] R. (Rolf) Karlsson and Andrzej Lingas, editors. Algorithm theory— SWAT ’96: 5th Scandinavian Workshop on Algorithm Theory, Reykjavik, Iceland, July 3– 5, 1996: proceedings, volume 1097 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61422-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .S28 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1097.htm;
 http://www.springerlink.
com/content/978-3-540-
61422-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1097.

Kuwabara:2002:IAM

[KL02] Kazuhiro Kuwabara and Jaeho Lee, editors. Intelligent Agents and Multi-Agent Systems: 5th Pacific Rim International Workshop on Multi Agents, Tokyo, Japan, August 19–20, 2002. Proceedings, volume 2413 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44026-7 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 P75 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2413.htm;
 http://www.springerlink.
com/content/978-3-540-
44026-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2413.

Kegl:2005:AAI

[KL05] Balázs Kégl and Guy Lapalme, editors. Advances in artificial intelligence: 18th conference of the Canadian Society for Computational Studies of Intelligence, Canadian AI 2005, Victoria, Canada, May 9–11, 2005, proceedings, volume 3501 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25864-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25864-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3501;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b136277.

Kao:2007:AAI

[KL07a] Ming-Yang Kao and Xiang-Yang Li, editors. Algorithmic Aspects in Information and Management: Third International Conference, AAIM 2007, Portland, OR, USA, June 6–8, 2007. Proceedings, volume 4508 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72868-6 (print), 3-540-72870-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72870-2.

Karssemeijer:2007:IPM

[KL07b] Nico Karssemeijer and Boudewijn Lelieveldt, editors. Information Processing in Medical Imaging: 20th International Conference, IPMI 2007, Kerkrade, The Netherlands, July 2–6, 2007. Proceedings, volume 4584 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73272-1 (print), 3-540-73273-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73273-0.

Kiayias:2012:VIT

[KL12] Aggelos Kiayias and Helger Lipmaa, editors. E-Voting and Identity: Third International Conference, VoteID 2011, Tallinn, Estonia, September 28– 30, 2011, Revised Selected Papers, volume 7187 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32746-X (print), 3-642-32747-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32747-6.

Klasing:2012:EAI

[Kla12] Ralf Klasing, editor. Experimental Algorithms: 11th International Symposium, SEA 2012, Bordeaux, France, June 7– 9, 2012. Proceedings, volume 7276 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30849-X (print), 3-642-30850-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30850-5.

Katsikas:2006:ISI

[KLB^{+}06] Sokratis K. Katsikas, Javier López, Michael Backes, Stefanos Gritzalis, and Bart Preneel, editors. Information Security: 9th International Conference, ISC 2006, Samos Island, Greece, August 30 — September 2, 2006. Proceedings, volume 4176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38341-7 (print), 3-540-38343-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        38343-7.

Kruegel:2007:RAI

[KLC07]  Christopher  Kruegel,

Richard Lippmann, and Andrew Clark, editors. Recent Advances in Intrusion Detection: 10th International Symposium, RAID 2007, Gold Goast, Australia, September 5–7, 2007. Proceedings, volume 4637 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74319-7 (print), 3-540-74320-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74320-0.

Klein:1989:CAV

[Kle89] Rolf Klein. Concrete and abstract Voronoi diagrams, volume 400 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-52055-4. ISSN 0302-9743 (print), 1611-3349 (electronic). iv + 167 pp. LCCN QA278.2 .K554 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0400.htm;
http://www.springerlink.
com/content/978-0-387-
52055-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=400.

Kratochvil:2010:TAM

[KLFK10] Jan Kratochvíl, Angsheng Li, Jiří Fiala, and Petr Kolman, editors.

Theory and Applications of Models of Computation: 7th Annual Conference, TAMC 2010, Prague, Czech Republic, June 7–11, 2010. Proceedings, volume 6108 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13561-7 (print), 3-642-13562-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13562-0.

Klint:1985:SSP

[Kli85] Paul Klint. A study in string processing languages, volume 205 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CODEN LNCSD9. ISBN 0-387-16041-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). various pp. LCCN QA76.9.T48 K581 1985. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0205.htm;
 http://www.springerlink.
com/content/978-0-387-
16041-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=205.

Cataloging based on CIP information.

Ko:2012:ADG

[KLLP12] Andrea Kő, Christine Leitner, Herbert Leitold, and Alexander Prosser, editors. Advancing Democracy, Government and Governance: Joint International Conference on Electronic Government and the Information Systems Perspective, and Electronic Democracy, EGOVIS/EDEM 2012, Vienna, Austria, September 3–6, 2012. Proceedings, volume 7452 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32700-1 (print), 3-642-32701-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32701-8.

Kedad:2007:NLP

[KLM^{+}07] Zoubida Kedad, Nadira Lammari, Elisabeth Métais, Farid Meziane, and Yacine Rezgui, editors. Natural Language Processing and Information Systems: 12th International Conference on Applications of Natural Language to Information Systems, NLDB 2007, Paris, France, June 27– 29, 2007. Proceedings, volume 4592 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73350-7 (print), 3-540-73351-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73351-5.

Kramer:2007:SOC

[KLN07] Bernd J. Krämer, Kwei-Jay Lin, and Priya Narasimhan, editors. Service-Oriented Computing — ICSOC 2007: Fifth International Conference, Vienna, Austria, September 17–20, 2007. Proceedings, volume 4749 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74973-X (print), 3-540-74974-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74974-5.

Kloos:1987:SDC

[Klo87] Carlos Delgado Kloos.

Semantics of digital circuits, volume 285 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18540-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 124 pp. LCCN TK7885.7 .K561 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0285.htm;
 http://www.springerlink.
com/content/978-0-387-
18540-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=285.

Kloks:1994:TCA

[Klo94] Ton Kloks. Treewidth:

computations and approximations, volume 842 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58356-4 (Berlin), 0-387-58356-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 209 pp. LCCN QA166.2 .K58 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0842.
htm;   http://www.springerlink.
com/content/978-3-540-
58356-1;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=842.

Katsikas:2004:TPD

[KLP04] Sokratis Katsikas, Javier Lopez, and Guenther Pernul, editors. Trust and Privacy in Digital Business: First International Conference, Trustbus 2004, Zaragoza, Spain, August 30–September 1, 2004, Proceedings, volume 3184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22919-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5978 2004.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3184.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        22919-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3184;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99832.

Khuri:2010:ITB

[KLP10] Sami Khuri, Lenka Lhotská, and Nadia Pisanti, editors. Information Technology in Bio- and Medical Informatics, IT-BAM 2010: First International Conference, Bilbao, Spain, September 1–2, 2010. Proceedings, volume 6266 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15019-5 (print), 3-642-15020-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15020-3.

Konstantas:2003:OOI

[KLPP03] Dimitri Konstantas, Michel Leonard, Yves Pigneur, and Shusma Patel, editors. Object-Oriented Information Systems: 9th International Conference, OOIS 2003: Geneva, Switzerland, September 2–5, 2003: Proceedings, volume 2817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40860-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55875 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2817.
htm;   http://www.springerlink.
com/content/978-3-540-

                                        40860-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2817;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13248.

Kaminka:2003:RRS

[KLR03] Gal A. Kaminka, Pedro U. Lima, and Raul Rojas, editors. RoboCup 2002: Robot Soccer World Cup VI, volume 2752 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40666-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ210.3 .R6153 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2752.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40666-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2752;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11927.

Katsikas:2010:TPS

[KLS10] Sokratis Katsikas, Javier Lopez, and Miguel Soriano, editors. Trust, Privacy and Security in Digital Business: 7th International Conference, TrustBus 2010, Bilbao, Spain, August 30–31, 2010. Proceedings, volume 6264 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15151-5 (print), 3-642-15152-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15152-1.

Kovacs:2007:LCS

[KLT^{+}07] Tim Kovacs, Xavier Llorà, Keiki Takadama, Pier Luca Lanzi, Wolfgang Stolzmann, and Stewart W. Wilson, editors. Learning Classifier Systems: International Workshops, IWLCS 2003-2005, Revised Selected Papers, volume 4399 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71230-5 (print), 3-540-71231-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        71231-2.

Kurz:2009:ACC

[KLT09] Alexander Kurz, Marina Lenisa, and Andrzej Tarlecki, editors. Algebra and Coalgebra in Computer Science: Third International Conference, CALCO 2009, Udine, Italy, September 7–10, 2009. Proceedings, volume 5728 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03740-2 (print), 3-642-03741-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03741-2.

Kay:2012:PCI

[KLT^{+}12] Judy Kay, Paul Lukowicz, Hideyuki Tokuda, Patrick Olivier, and Antonio Krüger, editors. Pervasive Computing: 10th International Conference, Pervasive 2012, Newcastle, UK, June 18– 22, 2012. Proceedings, volume 7319 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31204-7 (print), 3-642-31205-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31205-2.

Kluge:1997:IFL

[Klu97] Werner (Werner E.)

Kluge, editor. Implementation of functional languages: 8th International Workshop, Bad Godesberg, Germany, September 1996: Selected Papers, volume 1268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63237-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62.I58 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1268.
htm;   http://www.springerlink.
com/content/978-3-540-
63237-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1268.

Kang:2007:ACI

[KLZ07a] Lishan Kang, Yong Liu, and Sanyou Zeng, editors. Advances in Computation and Intelligence: Second International Symposium, ISICA 2007 Wuhan, China, September 21– 23, 2007 Proceedings, volume 4683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74580-7 (print), 3-540-74581-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74581-5.

Kang:2007:ESB

[KLZ07b] Lishan Kang, Yong Liu, and Sanyou Zeng, editors. Evolvable Systems: From Biology to Hardware: 7th International Conference, ICES 2007, Wuhan, China, September 21–23, 2007 Proceedings, volume 4684 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74625-0 (print), 3-540-74626-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74626-3.

Kim:1996:ACA

[KM96] Kwangjo Kim and Tsutomu Matsumoto, editors. Advances in cryptology—ASIACRYPT ’96: International Conference on the Theory and Applications of Cryptology and Information Security, Kyongju, Korea, November 3–7, 1996: proceedings, volume 1163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61872-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25I5553 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1163.htm;
http://www.springerlink.
com/content/978-3-540-
61872-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1163.

Kotz:2000:ASM

[KM00] David Kotz and Friedemann Mattern, editors. Agent systems, mobile agents, and applications: Second International Symposium on Agent Systems and Applications and Fourth International Symposium on Mobile Agents, ASA/MA 2000, Zürich, Switzerland, September 13–15, 2000: Proceedings, volume 1882 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41052-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1882. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1882.htm;
 http://www.springerlink.
com/content/978-3-540-
41052-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1882;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b75241.

King:2004:DDS

[KM04] Peter King and Ethan V. Munson, editors. Digital Documents: Systems and Principles: 8th International Conference on Digital Documents and Electronic Publishing, DDEP 2000 and 5th International Workshop on the Principles of Digital Document Processing, PODDP 2000, Munich, Germany, September 13–15, 2000: Revised Papers, volume 2023 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21070-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z286.E43 I565 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2023.htm;
 http://www.springerlink.
com/content/978-3-540-
21070-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2023;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95721.

Kaminski:2008:CSL

[KM08a] Michael Kaminski and Simone Martini, editors. Computer Science Logic: 22nd International Workshop, CSL 2008, 17th Annual Conference of the EACSL, Bertinoro, Italy, September 16–19, 2008. Proceedings, volume 5213 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87530-1 (print), 3-540-87531-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87531-4.

Kawano:2008:TQC

[KM08b] Yasuhito Kawano and Michele Mosca, editors. Theory of Quantum Computation, Communication, and Cryptography: Third Workshop, TQC 2008 Tokyo, Japan, January 30 — February 1, 2008. Revised Selected Papers, volume 5106 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89303-2 (print), 3-540-89304-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89304-2.

Kalas:2011:ISC

[KM11] Ivan Kalaš and Roland T. Mittermeir, editors. Informatics in Schools. Contributing to 21st Century Education: 5th International Conference on Informatics in Schools: Situation, Evolution and Perspectives, ISSEP 2011, Bratislava, Slovakia, October 26– 29, 2011. Proceedings, volume 7013 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24721-0 (print), 3-642-24722-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24722-4.

Katila:2001:FIM

[KMC^{+}01] Toivo E. Katila, Isabelle E. Magnin, Patrick Clarysse, Johan Montagnat, and Jukka Nenonen, editors. Functional Imaging and Modeling of the Heart: First International Workshop, Helsinki, Finland, November 15–16, 2001. Proceedings, volume 2230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42861-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Qp112.4 .I58 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2230.htm;
 http://www.springerlink.
com/content/978-3-540-
42861-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2230.

Karmouch:2002:MAT

[KMD02] Ahmed Karmouch, Thomas Magedanz, and Jaime Delgado, editors. Mobile Agents for Telecommunication Applications: 4th International Workshop, MATA 2002, Barcelona, Spain, October 23–24, 2002. Proceedings, volume 2521 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00021-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 I585 2002; TK5101.A1 .M319 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2521.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        00021-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2521.

Keller:2006:SMN

[KMF06] Alexander Keller and Jean-Philippe Martin-Flatin, editors. Self-Managed Networks, Systems, and Services: Second IEEE International Workshop, Self-Man 2006, Dublin, Ireland, June 16, 2006. Proceedings, volume 3996 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34739-9 (print), 3-540-34740-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34740-8.

Katz:2010:TAO

[KMK10] Shmuel Katz, Mira Mezini, and Jörg Kienzle, editors. Transactions on Aspect-Oriented Software Development VII: A Common Case Study for Aspect-Oriented Modeling, volume 6210 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16085-9 (print), 3-642-16086-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16086-8.

Kuznetsov:2011:PRM

[KMKP11] Sergei O. Kuznetsov, Deba P. Mandal, Malay K. Kundu, and Sankar K. Pal, editors. Pattern Recognition and Machine Intelligence: 4th International Conference, PReMI 2011, Moscow, Russia, June 27 — July 1, 2011. Proceedings, volume 6744 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21785-0 (print), 3-642-21786-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21786-9.

Kappel:2011:SOC

[KMMN11] Gerti Kappel, Zakaria Maamar, and Hamid R. Motahari-Nezhad, editors. Service-Oriented Computing: 9th International Conference, ICSOC 2011, Paphos, Cyprus, December 5– 8, 2011 Proceedings, volume 7084 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25534-5 (print), 3-642-25535-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25535-9.

Kundu:2012:PMI

[KMMP12] Malay K. Kundu, Sushmita Mitra, Debasis Mazumdar, and Sankar K. Pal, editors. Perception and Machine Intelligence: First Indo-Japan Conference, Per-MIn 2012, Kolkata, India, January 12–13, 2012. Proceedings, volume 7143 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27386-6 (print), 3-642-27387-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27387-2.

Koucheryavy:2012:WWI

[KMMT12] Yevgeni Koucheryavy, Lefteris Mamatas, Ibrahim Matta, and Vassilis Tsaoussidis, editors. Wired/ Wireless Internet Communication: 10th International Conference, WWIC 2012, Santorini, Greece, June 6–8, 2012. Proceedings, volume 7277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30629-2 (print), 3-642-30630-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30630-3.

Kugler:1994:TPE

[KMN94] Hans-Jurgen Kugler, Al Mullery, and Norbert Niebert, editors. Towards a Pan-European telecommunication service infrastructure: Second International Conference on Intelligence in Broadband Services and Networks, IS and N ’94, Aachen, Germany, September 7– 9, 1994: proceedings, volume 851 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58420-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5101.A1 I4798 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0851.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58420-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=851.

Kreowski:2005:FMS

[KMO^{+}05] Hans-Jörg Kreowski, Ugo Montanari, Fernando Orejas, Grzegorz Rozenberg, and Gabriele Taentzer, editors. Formal methods in software and systems modeling: Essays dedicated to Hartmut Ehrig on the occasion of his 60th birthday, volume 3393 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24936-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xxvii + 411 pp. LCCN QA76.9.F67 F66 2005.

URL http://www.springerlink.
com/content/978-3-540-
24936-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3393;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106390.

Kahn:1984:SDT

[KMP84] G. Kahn, D. B. Mac-Queen, and G. Plotkin, editors. Semantics of data types: International Symposium, Sophia-Antipolis, France, June 27–29, 1984: proceedings, volume 173 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-13346-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .S44 1984. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0173.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        13346-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=173.

“The symposium was sponsored by IFIP and EATCS”–Foreword. Proceedings of the International Symposium on the Semantics of Data Types. Cataloging based on CIP information.

Kipp:2009:MCM

[KMPH09] Michael Kipp, Jean-Claude Martin, Patrizia Paggio, and Dirk Heylen, editors. Multimodal Corpora: From Models of Natural Interaction to Systems and Applications, volume 5509 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04792-0 (print), 3-642-04793-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-642-
04793-0.

Karhumaki:2004:TFE

[KMPR04] Juhani Karhumäki, Hermann Maurer, Gheorghe Păun, and Grzegorz Rozenberg, editors. Theory Is Forever: Essays dedicated to Arto Salomaa on the occasion of his 70th birthday, volume 3113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22393-2. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 281 pp. LCCN QA76.9.M35 T44 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3113.htm;
http://www.springerlink.
com/content/978-3-540-
22393-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3113;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98751.

Kantor:2005:ISI

[KMR^{+}05] Paul Kantor, Gheorghe Muresan, Fred Roberts, Daniel D. Zeng, Fei-YueWang, Hsinchun Chen, and Ralph C. Merkle, editors. Intelligence and Security Informatics: IEEE International Conference on Intelligence and Security Informatics, ISI 2005, Atlanta, GA, USA, May 19–20, 2005, proceedings, volume 3495 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25999-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25999-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3495;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136511.

Kutrib:2012:DCF

[KMR12] Martin Kutrib, Nelma Moreira, and Rogério Reis, editors. Descriptional Complexity of Formal Systems: 14th International Workshop, DCFS 2012, Braga, Portugal, July 23–25, 2012. Proceedings, volume 7386 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31622-0 (print), 3-642-31623-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31623-4.

Kreczmar:1989:MFC

[KMS89] Antoni Kreczmar and Grazyna Mirkowska-Salwicka, editors. Mathematical foundations of computer science 1989: Porabka-Kozubnik, Poland, August 28–September 1, 1989: proceedings, volume 379 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51486-4 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 M386 1989.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0379.
htm;   http://www.springerlink.
com/content/978-0-387-
51486-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=379.

“Contains papers selected for presentation at the 14th Symposium on Mathematical Foundations of Computer Science, MFCS’89”–Pref.

Kohlas:2006:DSS

[KMS06] Jürg Kohlas, Bertrand Meyer, and André Schiper, editors. Dependable Systems: Software, Computing, Networks: Research Results of the DICS Program, volume 4028 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36821-3 (print), 3-540-36823-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        36823-6.

Kuri-Morales:2010:AAI

[KMS10] Angel Kuri-Morales and Guillermo R. Simari, editors. Advances in Artificial Intelligence — IB-ERAMIA 2010: 12th Ibero–American Conference on AI, Bahía Blanca, Argentina, November 1–5, 2010. Proceedings, volume 6433 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16951-1 (print), 3-642-16952-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16952-6.

Keyson:2011:AIS

[KMS^{+}11] David V. Keyson, Mary Lou Maher, Norbert Streitz, Adrian Cheok, Juan Carlos Augusto, Reiner Wichert, Gwenn Englebienne, Hamid Aghajan, and Ben J. A. Kröse, editors. Ambient Intelligence: Second International Joint Conference on AmI 2011, Amsterdam, The Netherlands, November 16–18, 2011. Proceedings, volume 7040 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25166-8 (print), 3-642-25167-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25167-2.

Kothe:2012:ADG

[KMS12] Ullrich Köthe, Annick Montanvert, and Pierre Soille, editors. Applications of Discrete Geometry and Mathematical Morphology: First International Workshop, WADGMM 2010, Istanbul, Turkey, August 22, 2010, Revised Selected Papers, volume 7346 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32312-X (print), 3-642-32313-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32313-3.

Kikuchi:2011:DNI

[KMSB11] Shinji Kikuchi, Aastha Madaan, Shelly Sachdeva, and Subhash Bhalla, editors. Databases in Networked Information Systems: 7th International Workshop, DNIS 2011, Aizu-Wakamatsu, Japan, December 12–14, 2011. Proceedings, volume 7108 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25730-5 (print), 3-642-25731-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25731-5.

Katz:2011:TAO

[KMSJ11] Shmuel Katz, Mira Mezini, Christine Schwanninger, and Wouter Joosen, editors. Transactions on Aspect-Oriented Software Development VIII, volume 6580 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22030-4 (print), 3-642-22031-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22031-9.

Kohavi:2002:WML

[KMSS02] Ron Kohavi, Brij Masand, Myra Spiliopoulou, and Jaideep Srivastava, editors. WEBKDD 2001— Mining Log Data Across All Customer Touch Points: Third International Workshop, San Francisco, CA, USA, August 26, 2001: Revised Papers, volume 2356 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43969-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4235 .W43 2001 Bar. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2356.

                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        43969-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2356.

Kambayashi:2000:DWK

[KMT00] Y. Kambayashi, Mukesh Mohania, and A. Min Tjoa, editors. Data Warehousing and Knowledge Discovery: Second International Conference, DaWaK 2000, London, UK, September 4–6, 2000. Proceedings, volume 1874 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67980-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D37 D396 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1874.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67980-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1874.

Kowalczyk:2003:ATI

[KMTU03] Ryszard Kowalczyk, J. P. Müller, H. Tianfield, and R. Unland, editors. Agent-technologies, infrastructures, tools, and applications for eservices: NODe 2002 agent-related workshops, Erfurt, Germany, October 7–19, 2002: Revised Papers, volume 2592 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00742-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 N63 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2592.htm;
 http://www.springerlink.
com/content/978-3-540-
00742-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2592.

Kalinichenko:2003:ADI

[KMTW03] L. A. (Leonid Andreevich) Kalinichenko, Rainer Manthey, Bernhard Thalheim, and Uwe Wloka, editors. Advances in Databases and Information Systems: 7th East European Conference, ADBIS 2003, Dresden, Germany, September 3–6, 2003: Proceedings, volume 2798 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20047-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2798.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20047-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2798;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12032.

Kronland-Martinet:2006:CMM

[KMVY06] Richard Kronland-Martinet, Thierry Voinier, and Sølvi Ystad, editors. Computer Music Modeling and Retrieval: Third International Symposium, CMMR 2005, Pisa, Italy, September 26–28, 2005. Revised Papers, volume 3902 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-34027-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34027-0;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3902.

Kambayashi:2003:DWK

[KMW03] Y. Kambayashi, Mukesh Mohania, and Wolfram Wöss, editors. Data Warehousing and Knowledge Discovery: 5th International Conference, DaWaK 2003 Prague, Czech Republic, September 3–5, 2003 Proceedings, volume 2737 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40807-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D37 D396 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2737.
htm;   http://www.springerlink.
com/content/978-3-540-
40807-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2737;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11825.

Kambayashi:2004:DWK

[KMW04] Yahiko Kambayashi, Mukesh Mohania, and Wolfram Woess, editors. Data Warehousing and Knowledge Discovery: 6th International Conference, DAWAK 2004, Zaragoza, Spain, September 1– 3, 2004: Proceedings, volume 3181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22937-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3181.htm;
 http://www.springerlink.
com/content/978-3-540-
22937-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3181;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99817.

Kronland-Martinet:2008:CMM

[KMYJ08] Richard Kronland-Martinet, Sølvi Ystad, and Kristoffer Jensen, editors. Computer Music Modeling and Retrieval. Sense of Sounds: 4th International Symposium, CMMR 2007, Copenhagen, Denmark, August 27–31, 2007. Revised Papers, volume 4969 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85034-1 (print), 3-540-85035-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85035-9.

Knuth:1983:SDS

[KN83] E. Knuth and Erich J.

Neuhold, editors. Specification and design of software systems: Conference on Operating Systems Visegrad, Hungary, January 23–27, 1982, proceedings, volume 152 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-12284-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .C646 1982. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0152.htm;
http://www.springerlink.
com/content/978-0-387-
12284-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=152.

Karshmer:1991:OSB

[KN91] A. Karshmer and J. Nehmer, editors. Operating systems of the 90s and beyond: international workshop, Dagstuhl Castle, Germany, July 8– 12, 1991: proceedings, volume 563 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-54987-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.O63 O64 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0563.
htm;   http://www.springerlink.
com/content/978-0-387-
54987-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=563.

Kittler:2003:AVB

[KN03] Josef Kittler and Mark S. Nixon, editors. Audioand video-based biometric person authentication: 4th International Conference, AVBPA 2003, Guildford, UK, June 9–11, 2003: Proceedings, volume 2688 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40302-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.S65 A944 2003. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t2688.htm;
http://www.springerlink.
com/content/978-3-540-
40302-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2688.

Kaklamanis:2009:TGC

[KN09a] Christos Kaklamanis and Flemming Nielson, editors. Trustworthy Global Computing: 4th International Symposium, TGC 2008, Barcelona, Spain, November 3–4, 2008, Revised Selected Papers, volume 5474 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00944-1 (print), 3-642-00945-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00945-7.

Kralovic:2009:MFC

[KN09b] Rastislav Královič and Damian Niwiński, editors. Mathematical Foundations of Computer Science 2009: 34th International Symposium, MFCS 2009, Novy Smokovec, High Tatras, Slovakia, August 24– 28, 2009. Proceedings, volume 5734 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03815-8 (print), 3-642-03816-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03816-7.

Krurkova:2008:ANNa

[KNK08a] Véra Kůrková, Roman Neruda, and Jan Koutník, editors. Artificial Neural Networks — ICANN 2008: 18th International Conference, Prague, Czech Republic, September 3–6, 2008, Proceedings, Part I, volume 5163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87535-2 (print), 3-540-87536-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87536-9.

Krurkova:2008:ANNb

[KNK08b] Véra Kůrková, Roman Neruda, and Jan Koutník, editors. Artificial Neural Networks — ICANN 2008: 18th International Conference, Prague, Czech Republic, September 3–6, 2008, Proceedings, Part II, volume 5164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87558-1 (print), 3-540-87559-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87559-8.

Knoop:1998:OIP

[Kno98] Jens Knoop. Optimal Interprocedural Program Optimization: A New Framework and Its Application, volume 1428 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65123-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xxv + 288 pp. LCCN QA76.6 .K63 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1428.
htm;   http://www.springerlink.
com/content/978-3-540-
65123-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1428.

Knoop:2011:CCI

[Kno11] Jens Knoop, editor. Compiler Construction: 20th International Conference, CC 2011, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2011, Saarbrücken, Germany, March 26–April 3, 2011. Proceedings, volume 6601 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19860-0 (print), 3-642-19861-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19861-8.

Koc:2001:CHEb

[KNP01]  Çetin K. Koç, David

Naccache, and Christof Paar, editors. Cryptographic hardware and embedded systems — CHES 2001: Third International Workshop, Paris, France, May 14– 16, 2001: Proceedings, volume 2162 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42521-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 C454 2001; QA267.A1 L43 no.2162.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2162.
htm;   http://www.springerlink.
com/content/978-3-540-
42521-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2162.

Kuba:2006:DGC

[KNP06] Attila Kuba, László G.

Nyúl, and Kálmán Palágyi, editors. Discrete Geometry for Computer Imagery: 13th International Conference, DGCI 2006, Szeged, Hungary, October 25–27, 2006. Proceedings, volume 4245 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47651-2 (print), 3-540-47652-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47652-8.

Knuth:1992:AH

[Knu92] Donald Ervin Knuth.

Axioms and hulls, volume 606 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55611-7 (Berlin), 0-387-55611-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 109 pp. LCCN QA640.3 .K68 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0606.
htm;   http://www.springerlink.
com/content/978-3-540-
55611-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=606.

Knudsen:1999:FSE

[Knu99a] Lars Knudsen, editor. Fast software encryption: 6th International Workshop, FSE’99, Rome, Italy, March 24–26, 1999: Proceedings, volume 1636 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66226-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F77 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1636.htm;

http://www.springerlink.
com/content/978-3-540-
66226-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1636.

Knuth:1999:MRC

[Knu99b] Donald Ervin Knuth.

MMIXware: a RISC computer for the third millennium, volume 1750 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. ISBN 3-540-66938-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 550 pp. LCCN QA76.9.A73 K62 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1750.
htm;   http://www.springerlink.
com/content/978-3-540-
66938-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1750.

Knudsen:2001:EOO

[Knu01] J. L. Knudsen, editor.

ECOOP 2001 — objectoriented programming: 15th European Conference, Budapest, Hungary, June 18–22, 2001: Proceedings, volume 2072 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42206-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 2001; QA267.A1 L43 no.2072. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2072.htm;
 http://www.springerlink.
com/content/978-3-540-
42206-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2072.

Knudsen:2002:ACE

[Knu02] Lars Knudsen, editor. Advances in Cryptology— EUROCRYPT 2002: International Conference on the Theory and Applications of Cryptographic Techniques, Amsterdam, The Netherlands, April 28–May 2, 2002. Proceedings, volume 2332 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43553-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2332.htm;
 http://www.springerlink.
com/content/978-3-540-
43553-2;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2332.

Kaplan:1991:CTR

[KO91] S. Kaplan and M. Okada, editors. Conditional and typed rewriting systems: 2nd International CTRS Workshop, Montréal, Canada, June 11–14, 1990: proceedings, volume 516 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54317-1 (Berlin), 0-387-54317-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .C5932 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0516.htm;
http://www.springerlink.
com/content/978-3-540-
54317-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=516.

Kranakis:2007:AHM

[KO07a] Evangelos Kranakis and Jaroslav Opatrny, editors. Ad-Hoc, Mobile, and Wireless Networks: 6th International Conference, ADHOC-NOW 2007, Morelia, Mexico, September 24–26, 2007, Proceedings, volume 4686 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74822-9 (print), 3-540-74823-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74823-6.

Krishnamurthi:2007:CCI

[KO07b] Shriram Krishnamurthi and Martin Odersky, editors. Compiler Construction: 16th International Conference, CC 2007, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2007, Braga, Portugal, March 26–30, 2007. Proceedings, volume 4420 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71228-3 (print), 3-540-71229-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71229-9.

Klugl:2011:MST

[KO11] Franziska Klügl and Sascha Ossowski, editors. Multiagent System Technologies: 9th German Conference, MATES 2011, Berlin, Germany, October 6–7, 2011. Proceedings, volume 6973 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24602-8 (print), 3-642-24603-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24603-6.

Koblitz:1996:ACC

[Kob96] Neal Koblitz, editor.

Advances in cryptology, CRYPTO ’96: 16th annual international cryptology conference, Santa Barbara, California, USA, August 18–22, 1996: proceedings, volume 1109 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61512-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1109.
htm;   http://www.springerlink.
com/content/978-3-540-
61512-5;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1109.

Sponsored by the International Association for Cryptologic Research (IACR), in cooperation with the IEEE Computer Society Technical Committee on Security and Privacy and the Computer Science Department of the University of California at Santa Barbara (UCSB).

Kobayashi:2006:PLS

[Kob06] Naoki Kobayashi, editor. Programming Languages and Systems: 4th Asian Symposium, APLAS 2006, Sydney, Australia, November 8– 10, 2006. Proceedings, volume 4279 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48937-1 (print), 3-540-48938-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        48938-2.

Kodratoff:1991:MLE

[Kod91] Yves Kodratoff, editor. Machine learning — EWSL-91: proceedings, volume 482 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-53816-X (Berlin), 0-387-53816-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E87 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0482.htm;
 http://www.springerlink.
com/content/978-3-540-
53816-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=482.

Katz:2009:TAO

[KOFJ09] Shmuel Katz, Harold Ossher, Robert France, and Jean-Marc Jézéquel, editors. Transactions on Aspect-Oriented Software Development VI: Special Issue on Aspects and Model-Driven Engineering, volume 5560 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03763-1 (print), 3-642-03764-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        03764-1.

Kohlhase:2006:MKM

[Koh06a] Michael Kohlhase, editor. Mathematical knowledge management: 4th international conference, MKM 2005, Bremen, Germany, July 15–17, 2005: revised selected papers, volume 3863 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31430-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .M56 2005; QA76.95 .I565 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        31430-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3863.

Kohlhase:2006:OOM

[Koh06b] Michael Kohlhase, editor. OMDoc — An Open Markup Format for Mathematical Documents [version 1.2]: Foreword by Allan Bundy, volume 4180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37897-9 (print), 3-540-37898-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 171 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37898-3.

Kojima:1991:UAI

[Koj91] M. Kojima. A Unified approach to interior point algorithms for linear complementarity problems, volume 538 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-54509-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 108 pp. LCCN QA402.5 .U55 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0538.htm;
http://www.springerlink.
com/content/978-0-387-
54509-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=538.

Spine title: Interior point algorithms for linear complementarity problems.

Klusch:2004:CIA

[KOK04] Matthias Klusch, Sascha Ossowski, and Vipul Kashyap, editors. Cooperative Information Agents VIII: 8th International Workshop, CIA 2004, Erfurt, Germany, September 27–29, 2004: Proceedings, volume 3191 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23170-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3191.htm;
 http://www.springerlink.
com/content/978-3-540-
23170-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3191;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100785.

Keijzer:2004:GPE

[KOL^{+}04] Maarten Keijzer, Una-May O’Reilly, Simon M. Lucas, Ernesto Costa, and Terence Soule, editors. Genetic Programming: 7th European Conference, EuroGP 2004, Coimbra, Portugal, April 5–7, 2004: Proceedings, volume 3003 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21346-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN xi + 410. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3003.htm;
 http://www.springerlink.
com/content/978-3-540-
21346-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3003;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96274.

Kompe:1997:PSU

[Kom97] Ralf Kompe. Prosody in speech understanding systems, volume 1307 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63580-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xix + 357 pp. LCCN TK7895.S65 K66 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1307.
htm;   http://www.springerlink.
com/content/978-3-540-
63580-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1307.

Konrad:2004:MGN

[Kon04] Karsten Konrad, editor. Model Generation for Natural Language Interpretation and Analysis, volume 2953 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21069-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 166 pp. LCCN P98 .K634 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2953.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21069-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2953;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95744.

Klusch:2003:CIA

[KOOL03] Matthias Klusch, Andrea Omicini, Sascha Ossowski, and Heimo Laamanen, editors. Cooperative Information Agents VII: 7th International Workshop, CIA 2003, Helsinki, Finland, August 27–29, 2003: Proceedings, volume 2782 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40798-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 C53 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2782.
htm;   http://www.springerlink.
com/content/978-3-540-
40798-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2782;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12011.

Koskimies:1998:CCI

[Kos98] Kai Koskimies, editor.

Compiler construction: 7th International Conference, CC ’98, held as part of the European Joint Conferences on the Theory and Practice of Software, ETAPS ’98, Lisboa, Portugal, March 28–April 4, 1998: Proceedings, volume 1383 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64304-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65I52 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1383.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64304-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1383.

Klusch:2002:CIA

[KOS02] Matthias Klusch, Sascha Ossowski, and Onn M. Shehory, editors. Cooperative Information Agents VI: 6th International Workshop, CIA 2002, Madrid, Spain, September 18–20, 2002. Proceedings, volume 2446 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44173-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q337 .C66 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2446.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        44173-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2446.

Krogstie:2007:AIS

[KOS07] John Krogstie, Andreas Opdahl, and Guttorm Sindre, editors. Advanced Information Systems Engineering: 19th International Conference, CAiSE 2007, Trondheim, Norway, June 11– 15, 2007. Proceedings, volume 4495 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72987-9 (print), 3-540-72988-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72988-4.

Kott:1986:ALP

[Kot86] Laurent Kott, editor.

Automata, languages, and programming: 13th international colloquium, Rennes, France, July 15–19, 1986: proceedings, volume 226 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16761-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.226. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0226.htm;

                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        16761-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=226.

ICALP (International Colloquium on Automata, Languages, and Programming) was sponsored by the European Association on Theoretical Computer Science and organized by INRIA.

Kouvatsos:2011:NPE

[Kou11] Demetres D. Kouvatsos, editor. Network Performance Engineering: A Handbook on Convergent Multi-Service Networks and Next Generation Internet, volume 5233 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-02741-5 (print), 3-642-02742-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02742-0.

Kowalczyk:2001:AAI

[Kow01] Ryszard Kowalczyk, editor. Advances in artificial intelligence: PRICAI 2000 workshop reader: four workshops held at PRICAI 2000, Melbourne, Australia, August 28- September 1, 2000: Revised Papers, volume 2112 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42597-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2112; Q334 .A347 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2112.htm;
 http://www.springerlink.
com/content/978-3-540-
42597-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2112.

Koymans:1992:SMP

[Koy92] R. (Ron) Koymans. Specifying message passing and time-critical systems with temporal logic, volume 651 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-56283-4 (Berlin), 0-387-56283-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 164 pp. LCCN TK5105 .K69 1992. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0651.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        56283-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=651.

Kozen:1982:LPW

[Koz82] Dexter Kozen, editor. Logics of Programs Workshop, Yorktown Heights, New York, May 1981, volume 131 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11212-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L594 1981. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0131.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        11212-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=131.

Kozen:2004:MPC

[Koz04] Dexter Kozen, editor. Mathematics of Program Construction: 7th International Conference, MPC 2004, Stirling, Scotland, UK, July 12– 14, 2004: Proceedings, volume 3125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22380-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .M3661 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3125.htm;
 http://www.springerlink.
com/content/978-3-540-
22380-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3125;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98756.

Kastens:1992:CCI

[KP92] Uwe Kastens and P. Pfahler, editors. Compiler construction: 4th International Conference CC’92, Paderborn, FRG, October 5–7, 1992: proceedings, volume 641 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55984-1, 0-387-55984-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 1992.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t0641.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        55984-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=641.

Koc:1999:CHE

[KP99]  Çetin K. Koç and Christof

Paar, editors. Cryptographic hardware and embedded systems: First International Workshop, CHES ’99, Worchester, MA, USA, August 1999: Proceedings, volume 1717 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66646-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1717.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1717.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66646-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1717.

Kapoor:2000:FTF

[KP00] Sanjiv Kapoor and Sanjiva Prasad, editors. FST TCS 2000: Foundations of Software Technology and Theoretical Computer Science: 20th Conference, New Delhi, India, December 13–15, 2000. Proceedings, volume 1974 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41413-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1974. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1974.htm;
 http://www.springerlink.
com/content/978-3-540-
41413-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1974.

Kobayashi:2001:TAC

[KP01a] Naoki Kobayashi and Benjamin C. Pierce, editors. Theoretical aspects of computer software: 4th International Symposium, TACS 2001, Sendai, Japan, October 29–31, 2001: Proceedings, volume 2215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42736-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2215. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2215.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42736-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2215.

Koc:2001:CHEa

[KP01b]  Çetin K. Koç and Christof

Paar, editors. Cryptographic hardware and embedded systems — CHES 2000: Second International Workshop, Worcester, MA, USA, August 17–18, 2000: Proceedings, volume 1965 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41455-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 C454 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1965.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41455-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1965.

Kazman:2004:CBS

[KP04] Rick Kazman and Dan Port, editors. COTS-Based Software Systems: Third International Conference, IC-CBSS 2004, Redondo Beach, CA, USA, February 1–4, 2004: Proceedings, volume 2959 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21903-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .I25 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2959.htm;
 http://www.springerlink.
com/content/978-3-540-
21903-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2959;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96987.

Kramer:2005:ILP

[KP05] Stefan Kramer and Bernhard Pfahringer, editors. Inductive Logic Programming: 15th International Conference, ILP 2005, Bonn, Germany, August 10–13, 2005. Proceedings, volume 3625 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28177-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28177-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3625.

Kalra:2006:CVG

[KP06] Prem K. Kalra and Shmuel Peleg, editors. Computer Vision, Graphics and Image Processing: 5th Indian Conference, ICVGIP 2006, Madurai, India, December 13–16, 2006. Proceedings, volume 4338 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68301-1 (print), 3-540-68302-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68302-5.

Kowalewski:2009:TAC

[KP09] Stefan Kowalewski and Anna Philippou, editors. Tools and Algorithms for the Construction and Analysis of Systems: 15th International Conference, TACAS 2009, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2009, York, UK, March 22– 29, 2009. Proceedings, volume 5505 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00767-8 (print), 3-642-00768-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00768-2.

Kaufmann:2010:ITP

[KP10a] Matt Kaufmann and Lawrence C. Paulson, editors. Interactive Theorem Proving: First International Conference, ITP 2010, Edinburgh, UK, July 11–14, 2010. Proceedings, volume 6172 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14051-3 (print), 3-642-14052-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14052-5.

Krishnamurthy:2010:PAM

[KP10b] Arvind Krishnamurthy and Bernhard Plattner, editors. Passive and Active Measurement: 11th International Conference, PAM 2010, Zurich, Switzerland, April 7– 9, 2010. Proceedings, volume 6032 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12333-3 (print), 3-642-12334-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12334-4.

Kucera:2010:RPI

[KP10c] Antonín Kučera and Igor Potapov, editors. Reachability Problems: 4th International Workshop, RP 2010, Brno, Czech Republic, August 28– 29, 2010. Proceedings, volume 6227 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15348-8 (print), 3-642-15349-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15349-5.

Kristensen:2011:ATP

[KP11] Lars M. Kristensen and Laure Petrucci, editors. Applications and Theory of Petri Nets: 32nd International Conference, PETRI NETS 2011, Newcastle, UK, June 20–24, 2011. Proceedings, volume 6709 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21833-4 (print), 3-642-21834-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21834-7.

Kroening:2015:CAV

[KP15] Daniel Kroening and Corina S. Păsăreanu, editors. Computer Aided Verification: 27th International Conference, CAV 2015, San Francisco, CA, USA, July 18– 24, 2015, Proceedings, Part I, volume 9206 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2015. ISBN 3-319-21689-9. URL http:

//link.springer.com/
book/10.1007/978-3-
319-21690-4.

Kalviainen:2005:IAS

[KPK05] Heikki Kalviainen, Jussi Parkkinen, and Arto Kaarna, editors. Image Analysis: 14th Scandinavian Conference, SCIA 2005, Joensuu, Finland, June 19–22, 2005. Proceedings, volume 3540 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26320-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .S29 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26320-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3540.

Konstantopoulos:2010:AIT

[KPK^{+}10] Stasinos Konstantopoulos, Stavros Perantonis, Vangelis Karkaletsis, Constantine D. Spyropoulos, and George Vouros, editors. Artificial Intelligence: Theories, Models and Applications: 6th Hellenic Conference on AI, SETN 2010, Athens, Greece, May 4–7, 2010. Proceedings, volume 6040 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12841-6 (print), 3-642-12842-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12842-4.

Klusch:2008:CIA

[KPP08] Matthias Klusch, Michal Pěchouček, and Axel Polleres, editors. Cooperative Information Agents XII: 12th International Workshop, CIA 2008, Prague, Czech Republic, September 10–12, 2008. Proceedings, volume 5180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85833-4 (print), 3-540-85834-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85834-8.

Krishnaswamy:2007:RTM

[KPR07] Dilip Krishnaswamy, Tom Pfeifer, and Danny Raz, editors. Real-Time Mobile Multimedia Services: 10th IFIP/IEEE International Conference on Management of Multimedia and Mobile Networks and Services, MMNS 2007, San José, USA, October 31 — November 2, 2007 Proceedings, volume 4787 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75868-2 (print), 3-540-75869-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75869-3.

Kryszkiewicz:2007:RSI

[KPRS07] Marzena Kryszkiewicz, James F. Peters, Henryk Rybinski, and Andrzej Skowron, editors. Rough Sets and Intelligent Systems Paradigms: International Conference, RSEISP 2007, Warsaw, Poland, June 28– 30, 2007. Proceedings, volume 4585 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73450-3 (print), 3-540-73451-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73451-2.

Klette:2001:RVI

[KPS01] Reinhard Klette, Shmuel Peleg, and Gerald Sommer, editors. Robot Vision: International Workshop RobVis 2001, Auckland, New Zealand, February 16–18, 2001: Proceedings, volume 1998 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41694-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ211.3 .I58 2001; QA267.A1 L43 no.1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1998.htm;
http://www.springerlink.
com/content/978-3-540-
41694-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1998.

Koopman:2009:AFP

[KPS09] Pieter Koopman, Rinus Plasmeijer, and Doaitse Swierstra, editors. Advanced Functional Programming: 6th International School, AFP 2008, Heijen, The Netherlands, May 2008, Revised Lectures, volume 5832 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04651-7 (print), 3-642-04652-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04652-0.

Kaklamanis:2012:EPP

[KPS12] Christos Kaklamanis, Theodore Papatheodorou, and Paul G. Spirakis, editors. Euro-Par 2012 Parallel Processing: 18th International Conference, Euro-Par 2012, Rhodes Island, Greece, August 27–31, 2012. Proceedings, volume 7484 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32819-9 (print), 3-642-32820-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32820-6.

Kant:2010:DCN

[KPSW10] Krishna Kant, Sriram V. Pemmaraju, Krishna M. Sivalingam, and Jie Wu, editors. Distributed Computing and Networking: 11th International Conference, ICDCN 2010, Kolkata, India, January 3–6, 2010. Proceedings, volume 5935 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11321-4 (print), 3-642-11322-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11322-2.

Kutylowski:1999:MFC

[KPW99] Miroslaw Kutylowski, Leszek Pacholski, and Tomasz Wierzbicki, editors. Mathematical foundations of computer science 1999: 24th International Symposium, MFCS’99, Szklarska Poreba, Poland, September 6–10, 1999: Proceedings, volume 1672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66408-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1672; QA76.9.M35 S99 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1672.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        66408-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1672.

Kirchner:2000:FCS

[KR00a] H. Kirchner and Christophe Ringeissen, editors. Frontiers of combining systems: Third International Workshop, FroCoS 2000, Nancy, France, March 22–24, 2000: Proceedings, volume 1794 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67281-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1794; QA9.A1 F73 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1794.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67281-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1794.

Kittler:2000:MCS

[KR00b] Josef Kittler and Fabio Roli, editors. Multiple classifier systems: First International Workshop, MCS 2000, Cagliari, Italy, June 21–23, 2000: Proceedings, volume 1857 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67704-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1857; Q325.5 .M84 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1857.htm;
 http://www.springerlink.
com/content/978-3-540-
67704-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1857.

Kittler:2001:MCS

[KR01] Josef Kittler and Fabio Roli, editors. Multiple classifier systems: Second International Workshop, MCS 2001, Cambridge, UK, July 2– 4, 2001: Proceedings, volume 2096 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42284-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2096; Q325.5 .M84 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2096.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42284-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2096.

Karagiannis:2002:PAK

[KR02a] D. Karagiannis and Ulrich Reimer, editors. Practical Aspects of Knowledge Management: 4th International Conference, PAKM 2002 Vienna, Austria, December 2–3, 2002. Proceedings, volume 2569 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00314-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HD30.2 .P353 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2569.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00314-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2569.

Also available via the

World Wide Web.

Kirchner:2002:AMS

[KR02b] H. Kirchner and C. Ringeissen, editors. Algebraic Methodology and Software Technology: 9th International Conference, AMAST 2002, Saint-Gilles-les-Bains, Reunion Island, France, September 9–13, 2002. Proceedings, volume 2422 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44144-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 I5713 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2422.htm;
http://www.springerlink.
com/content/978-3-540-
44144-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2422.

Krishnamurthi:2002:PAD

[KR02c] Shriram Krishnamurthi and C. R. Ramakrishnan, editors. Practical Aspects of Declarative Languages: 4th International Symposium, PADL 2002, Portland, OR, USA, January 19– 20, 2002: Proceedings, volume 2257 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43092-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .P32 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2257.htm;
 http://www.springerlink.
com/content/978-3-540-
43092-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2257.

Karagiannis:2004:PAK

[KR04a]  Dimitris  Karagiannis

and Ulrich Reimer, editors. Practical Aspects of Knowledge Management: 5th International Conference, PAKM 2004, Vienna, Austria, December 2–3, 2004: Proceedings, volume 3336 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24088-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HD30.2 .P353 2004. URL http:

//www.springerlink.
com/content/978-3-540-
24088-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3336;

http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104042.

Kuijpers:2004:CDA

[KR04b] Bart Kuijpers and Peter Revesz, editors. Constraint Databases and Applications: First International Symposium, CDB 2004, Paris, France, June 12–13, 2004: Proceedings, volume 3074 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22126-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C618 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3074.htm;
http://www.springerlink.
com/content/978-3-540-
22126-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3074;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98003.

Kunz:2006:AHM

[KR06] Thomas Kunz and S. S.

Ravi, editors. Ad-Hoc, Mobile, and Wireless Networks: 5th International Conference, ADHOC-NOW 2006, Ottawa, Canada, August 17–19, 2006. Proceedings, volume 4104 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37246-6 (print), 3-540-37248-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37248-6.

Kang:2010:KMA

[KR10a] Byeong-Ho Kang and Debbie Richards, editors. Knowledge Management and Acquisition for Smart Systems and Services: 11th International Workshop, PKAW 2010, Daegue, Korea, August 20 — September 3, 2010. Proceedings, volume 6232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15036-5 (print), 3-642-15037-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15037-1.

Kowalewski:2010:FMI

[KR10b] Stefan Kowalewski and Marco Roveri, editors. Formal Methods for Industrial Critical Systems: 15th International Workshop, FMICS 2010, Antwerp, Belgium, September 20–21, 2010. Proceedings, volume 6371 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15897-8 (print), 3-642-15898-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15898-8.

Kuich:2011:AFC

[KR11] Werner Kuich and George Rahonis, editors. Algebraic Foundations in Computer Science: Essays Dedicated to Symeon Bozapalidis on the Occasion of His Retirement, volume 7020 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24896-9 (print), 3-642-24897-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24897-9.

Kuncak:2012:VMC

[KR12] Viktor Kuncak and Andrey Rybalchenko, editors. Verification, Model Checking, and Abstract Interpretation: 13th International Conference, VMCAI 2012, Philadelphia, PA, USA, January 22–24, 2012. Proceedings, volume 7148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27939-2 (print), 3-642-27940-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27940-9.

Kraetzschmar:1997:DRM

[Kra97] Gerhard Kraetzschmar. Distributed reason maintenance for multiagent systems, volume 1229 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63606-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 296 pp. LCCN QA76.76.I58 K73 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1229.
htm;   http://www.springerlink.
com/content/978-3-540-
63606-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1229.

Krawczyk:1998:ACC

[Kra98] Hugo Krawczyk, editor. Advances in cryptology — CRYPTO ’98: 18th Annual International Cryptology Conference, Santa Barbara, California, USA August 23–27, 1998 proceedings, volume 1462 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64892-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1462.htm;
 http://www.springerlink.
com/content/978-3-540-
64892-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1462.

Sponsored by the International Association for Cryptologic Reseach, (IACR), in cooperation with the IEEE Computer Society Technical Committee on Security and Privacy and the Computer Science Department, University of California, Santa Barbara (UCSB).

Kratochvil:1999:GDI

[Kra99] Jan Kratochvil, editor. Graph drawing: 7th International Symposium, GD’99, Stirin Castle, Czech Republic, September 15–19, 1999: Proceedings, volume 1731 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66904-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .G69344 1999; T385 .S884 2000.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1731.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66904-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1731.

Krall:2003:SCE

[Kra03] Andreas Krall, editor. Software and Compilers for Embedded Systems: 7th International Workshop, SCOPES 2003, Vienna, Austria, September 24–26, 003: Proceedings, volume 2826 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20145-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2826.htm;
 http://www.springerlink.
com/content/978-3-540-
20145-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2826;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13482.

Kratsch:2005:GTC

[Kra05] Dieter (Dieter) Kratsch, editor. Graph-theoretic concepts in computer science: 31st international workshop, WG 2005, Metz, France, June 23– 25, 2005: revised selected papers, volume 3787 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-31000-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .I572 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        31000-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3787.

Korb:2009:ALB

[KRH09] Kevin Korb, Marcus Randall, and Tim Hendtlass, editors. Artificial Life: Borrowing from Biology: 4th Australian Conference, ACAL 2009, Melbourne, Australia, December 1–4, 2009. Proceedings, volume 5865 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10426-6 (print), 3-642-10427-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10427-5.

Konig-Ries:2002:DIM

[KRMM^{+}02] Birgitta König-Ries, Kia Makki, S. A. M. Makki, Niki Pissinou, and Peter Scheuermann, editors. Developing an infrastructure for mobile and wireless systems: NSF Workshop IMWS 2001, Scottsdale, AZ, October 15, 2001: Revised Papers, volume 2538 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00289-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK6570.M6 I45 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2538.
htm;   http://www.springerlink.
com/content/978-3-540-
00289-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2538.

Also available via the World Wide Web with additional title: Developing an infrastructure for mobile and wireless systems, NSF workshop IMWS 2001, Scottsdale, AZ, October 15, 2002: Revised Papers.

Kropf:2000:DCW

[Kro00] Peter Kropf, editor. Distributed communities on the Web: Third International Workshop, DCW 2000, Quebec City, Canada, June 19–21, 2000: Proceedings, volume 1830 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67647-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D375 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1830.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67647-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1830.

Klusch:2006:CIA

[KRP06] Matthias Klusch, Michael Rovatsos, and Terry R. Payne, editors. Cooperative Information Agents X: 10th International Workshop, CIA 2006 Edinburgh, UK, September 11–13, 2006 Proceedings, volume 4149 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38569-X (print), 3-540-38570-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        38570-7.

Kokinov:2007:MUC

[KRRBV07] Boicho Kokinov, Daniel C. Richardson, Thomas R. Roth-Berghofer, and Laure Vieu, editors. Modeling and Using Context: 6th International and Interdisciplinary Conference, CONTEXT 2007, Roskilde, Denmark, August 20– 24, 2007. Proceedings, volume 4635 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74254-9 (print), 3-540-74255-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74255-5.

Kuich:2002:DLT

[KRS02] Werner Kuich, Grzegorz Rozenberg, and Arto Salomaa, editors. Developments in language theory: 5th International Conference, DLT 2001, Wien, Austria, July 16– 21, 2001: Revised Papers, volume 2295 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43453-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.3 .C63 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2295.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        43453-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2295.

Kryszkiewicz:2011:FIS

[KRSR11] Marzena Kryszkiewicz, Henryk Rybinski, Andrzej Skowron, and Zbigniew W. Raś, editors. Foundations of Intelligent Systems: 19th International Symposium, ISMIS 2011, Warsaw, Poland, June 28–30, 2011. Proceedings, volume 6804 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21915-2 (print), 3-642-21916-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21916-0.

Krupinski:2008:DMI

[Kru08] Elizabeth A. Krupinski, editor. Digital Mammography: 9th International Workshop, IWDM 2008 Tucson, AZ, USA, July 20–23, 2008 Proceedings, volume 5116 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70537-6 (print), 3-540-70538-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70538-3.

Kruse:1991:SQA

[KS91] Rudolf Kruse and P. Siegel, editors. Symbolic and quantitative approaches to uncertainty: European Conference ECSQAU, Marseille, France, October 15–17, 1991: proceedings, volume 548 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54659-6 (Berlin), 0-387-54659-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q375 .E87 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0548.
htm;   http://www.springerlink.
com/content/978-3-540-
54659-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=548.

Klement:1993:FLA

[KS93] E. P. Klement and W. Slany, editors. Fuzzy logic in artificial intelligence: 8th Austrian Artificial Intelligence Conference, FLAI ’93, Linz, Austria, June 28–30, 1993: proceedings, volume 695 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56920-0 (Berlin), 0-387-56920-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q335 .A97 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0695.
htm;   http://www.springerlink.
com/content/978-3-540-
56920-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=695.

Klas:1995:MTA

[KS95] Wolfgang Klas and Michael Schrefl. Metaclasses and their application: data model tailoring and database integration, volume 943 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60063-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 200 pp. LCCN QA267.A1 L43 no.943.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0943.
htm;   http://www.springerlink.
com/content/978-3-540-
60063-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=943.

Kuchen:1996:PLI

[KS96] Herbert R. Kuchen and S. Doaitse Swierstra, editors. Programming languages: implementations, logics, and programs: 8th International Symposium, PLILP ’96, Aachen, Germany, September 24–27, 1996: proceedings, volume 1140 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61756-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7.P56 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1140.
htm;   http://www.springerlink.
com/content/978-3-540-
61756-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1140.

Kropf:1997:FHV

[KS97] Thomas Kropf and Mark Schneider. Formal hardware verification: methods and systems in comparison, volume 1287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63475-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 367 pp. LCCN TK7874.75.F67 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1287.
htm;   http://www.springerlink.
com/content/978-3-540-
63475-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1287.

Krakowiak:2000:ADS

[KS00] Sacha Krakowiak and S. K. Shrivastava, editors. Advances in Distributed Systems: Advanced Distributed Computing: From Algorithms to Systems, volume 1752 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67196-X. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 508 pp. LCCN QA76.9.D5 A363 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1752.htm;
 http://www.springerlink.

com/content/978-3-540-
67196-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1752.

Kelemen:2001:AAL

[KS01] Jozef Kelemen and Petr Sosik, editors. Advances in artificial life: 6th European Conference, ECAL 2001, Prague, Czech Republic, September 10–14, 2001: Proceedings, volume 2159 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42567-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .E87 2001; QA267.A1 L43 no.2159. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2159.htm;
http://www.springerlink.
com/content/978-3-540-
42567-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2159.

Kakas:2002:CLLa

[KS02a] Antonis C. Kakas and Fariba Sadri, editors. Computational Logic: Logic Programming and Beyond: Essays in Honour of Robert A. Kowalski, Part I, volume 2407 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43959-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 678 pp. LCCN QA76.63 .C64 2002. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2407.
htm;   http://www.springerlink.
com/content/978-3-540-
43959-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2407.

Kakas:2002:CLLb

[KS02b] Antonis C. Kakas and Fariba Sadri, editors. Computational Logic: Logic Programming and Beyond: Essays in Honour of Robert A. Kowalski, Part II, volume 2408 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43960-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 628 pp. LCCN QA76.63 .C64 2002. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2408.

htm;   http://www.springerlink.
com/content/978-3-540-
43960-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2408.

Katoen:2002:TAC

[KS02c] Joost-Pieter Katoen and Perdita Stevens, editors. Tools and algorithms for the construction and analysis of systems: 8th International Conference, TACAS 2002, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2002, Grenoble, France, April 8–12, 2002: Proceedings, volume 2280 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43419-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T33 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2280.htm;
 http://www.springerlink.
com/content/978-3-540-
43419-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2280.

Kivinen:2002:CLT

[KS02d] Jyrki Kivinen and Robert H. Sloan, editors. Computational Learning Theory: 15th Annual Conference on Computational Learning Theory, COLT 2002, Sydney, Australia, July 8–10, 2002. Proceedings, volume 2375 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43836-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .C66 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2375.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43836-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2375.

Kaashoek:2003:PPS

[KS03a] Frans Kaashoek and Ion Stoica, editors. Peer-to-Peer Systems II: Second International Workshop, IPTPS 2003 Berkeley, CA, USA, February 21– 22, 2003 Revised Papers, volume 2735 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40724-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.525 .I68 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2735.
htm;   http://www.springerlink.
com/content/978-3-540-
40724-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2735;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11823.

Karlsson:2003:QAC

[KS03b] Gunnar Karlsson and Michael I. Smirnov, editors. Quality for All: 4th COST 263 International Workshop on Quality of Future Internet Services, QoFIS 2003, Stockholm, Sweden, October 1–2, 2003: Proceedings, volume 2811 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20192-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 C67 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2811.

                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20192-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2811;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13241.

Kemper:2003:CPM

[KS03c] Peter Kemper and William Huston Sanders, editors. Computer Performance: Modelling techniques and tools: 13th International Conference, TOOLS 2003, Urbana, IL, USA, September 2–5, 2003: Proceedings, volume 2794 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40814-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.E94 T66 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2794.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40814-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2794;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12028.

Kameyama:2004:FLP

[KS04a] Yukiyoshi Kameyama and Peter J. Stuckey, editors. Functional and Logic Programming: 7th International Symposium, FLOPS 2004, Nara, Japan, April 7– 9, 2004: Proceedings, volume 2998 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21402-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .F84 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2998.htm;
http://www.springerlink.
com/content/978-3-540-
21402-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2998;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96926.

Koch:2004:RAT

[KS04b] Traugott Koch and Ingeborg Torvik Sølvberg, editors. Research and Advanced Technology for Digital Libraries: 7th European Conference, ECDL 2003, Trondheim, Norway, August 17–22, 2003: Proceedings, volume 2769 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-40726-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4080 .E28 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2769.htm;
 http://www.springerlink.
com/content/978-3-540-
40726-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2769;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11967.

Kralovic:2004:SIC

[KS04c] Rastislav Královič and Ondrej Sýkora, editors. Structural Information and Communication Complexity: 11th international colloquium, SIROCCO 2004, Smolenice Castle, Slowakia, June 21– 23, 2004: Proceedings, volume 3104 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22230-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 C64 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3104.
htm;   http://www.springerlink.
com/content/978-3-540-
22230-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3104;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98251.

Konagaya:2005:GCL

[KS05a] Akihiko Konagaya and Kenji Satou, editors. Grid computing in life science: first International Workshop on Life Science Grid, LSGRID 2004, Kanazawa, Japan, May 31–June 1, 2004: Revised selected and invited papers, volume 3370 of Lecture Notes in Bioinformatics and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25208-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 G75 2004. URL http:

//link.springer.com/
book/10.1007/b106923/
;   http://springerlink.
metapress.com/link.
asp?id=v1pnbm26b2m3;
 http://springerlink.

                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3370;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        25208-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3370;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106923.

Kotsis:2005:WSM

[KS05b] Gabriele Kotsis and Otto Spaniol, editors. Wireless systems and mobility in next generation internet: first international workshop of EURO-NGI network of excellence, Dagstuhl Castle, Germany, June 7– 9, 2004. Revised selected papers, volume 3427 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25329-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.2 .I59 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25329-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3427;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=

volume&id=doi:10.1007/
b107131.

Kim:2006:GMP

[KS06] Myung-Soo Kim and Kenji Shimada, editors. Geometric Modeling and Processing — GMP 2006: 4th International Conference, Pittsburgh, PA, USA, July 26–28, 2006. Proceedings, volume 4077 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36711-X (print), 3-540-36865-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36865-6.

Kordon:2007:CES

[KS07a] Fabrice Kordon and Oleg Sokolsky, editors. Composition of Embedded Systems. Scientific and Industrial Issues: 13th Monterey Workshop 2006 Paris, France, October 16–18, 2006 Revised Selected Papers, volume 4888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77418-1 (print), 3-540-77419-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77419-8.

Kordon:2007:RSU

[KS07b] Fabrice Kordon and Janos Sztipanovits, editors. Reliable Systems on Unreliable Networked Platforms: 12th Monterey Workshop 2005, Laguna Beach, CA, USA, September 22–24, 2005. Revised Selected Papers, volume 4322 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71155-4 (print), 3-540-71156-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71156-8.

Kuznetsov:2007:FCA

[KS07c] Sergei O. Kuznetsov and Stefan Schmidt, editors. Formal Concept Analysis: 5th International Conference, ICFCA 2007, Clermont-Ferrand, France, February 12–16, 2007. Proceedings, volume 4390 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70828-6 (print), 3-540-70901-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70901-5.

Kaklamanis:2008:AOA

[KS08a] Christos Kaklamanis and Martin Skutella, editors. Approximation and Online Algorithms: 5th International Workshop, WAOA 2007, Eilat, Israel, October 11–12, 2007. Revised Papers, volume 4927 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77917-5 (print), 3-540-77918-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77918-6.

Kulkarni:2008:SSS

[KS08b] Sandeep Kulkarni and André Schiper, editors. Stabilization, Safety, and Security of Distributed Systems: 10th International Symposium, SSS 2008, Detroit, MI, USA, November 21–23, 2008. Proceedings, volume 5340 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89334-2 (print), 3-540-89335-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89335-6.

Kaeli:2009:CPE

[KS09a] David Kaeli and Kai Sachs, editors. Computer Performance Evaluation and Benchmarking: SPEC Benchmark Workshop 2009, Austin, TX, USA, January 25, 2009. Proceedings, volume 5419 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-93798-6 (print), 3-540-93799-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        93799-9.

Katzenbeisser:2009:IHI

[KS09b] Stefan Katzenbeisser and Ahmad-Reza Sadeghi, editors. Information Hiding: 11th International Workshop, IH 2009, Darmstadt, Germany, June 8–10, 2009, Revised Selected Papers, volume 5806 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04430-1 (print), 3-642-04431-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04431-1.

Kotenko:2010:CNS

[KS10a] Igor Kotenko and Victor Skormin, editors. Computer Network Security: 5th International Conference on Mathematical Methods, Models and Architectures for Computer Network Security, MMM-ACNS 2010, St. Petersburg, Russia, September 8–10, 2010. Proceedings, volume 6258 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14705-4 (print), 3-642-14706-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14706-7.

Kumar:2010:AMW

[KS10b] Ravi Kumar and Dandapani Sivakumar, editors. Algorithms and Models for the Web-Graph: 7th International Workshop, WAW 2010, Stanford, CA, USA, December 13–14, 2010. Proceedings, volume 6516 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-18008-6 (print), 3-642-18009-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18009-5.

Kwuida:2010:FCA

[KS10c] Léonard Kwuida and Barış Sertkaya, editors. Formal Concept Analysis: 8th International Conference, ICFCA 2010, Agadir, Morocco, March 15– 18, 2010. Proceedings, volume 5986 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11927-1 (print), 3-642-11928-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11928-6.

Karkkainen:2012:CPM

[KS12a] Juha Kärkkäinen and Jens Stoye, editors. Combinatorial Pattern Matching: 23rd Annual Symposium, CPM 2012, Helsinki, Finland, July 3–5, 2012. Proceedings, volume 7354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31264-0 (print), 3-642-31265-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31265-6.

Khurshid:2012:RVS

[KS12b] Sarfraz Khurshid and Koushik Sen, editors. Runtime Verification: Second International Conference, RV 2011, San Francisco, CA, USA, September 27–30, 2011, Revised Selected Papers, volume 7186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29859-1 (print), 3-642-29860-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29860-8.

Kotenko:2012:CNS

[KS12c] Igor Kotenko and Victor Skormin, editors. Computer Network Security: 6th International Conference on Mathematical Methods, Models and Architectures for Computer Network Security, MMM-ACNS 2012, St. Petersburg, Russia, October 17–19, 2012. Proceedings, volume 7531 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33703-1 (print), 3-642-33704-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33704-8.

Krotzsch:2012:WRR

[KS12d] Markus Krötzsch and Umberto Straccia, editors. Web Reasoning and Rule Systems: 6th International Conference, RR 2012, Vienna, Austria, September 10–12, 2012. Proceedings, volume 7497 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33202-1 (print), 3-642-33203-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33203-6.

Kikuchi:2010:DNI

[KSB10] Shinji Kikuchi, Shelly Sachdeva, and Subhash Bhalla, editors. Databases in Networked Information Systems: 6th International Workshop, DNIS 2010, Aizu-Wakamatsu, Japan, March 29–31, 2010. Proceedings, volume 5999 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12037-7 (print), 3-642-12038-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12038-1.

Kim:2011:DWI

[KSB11] Hyoung-Joong Kim, Yun Qing Shi, and Mauro Barni, editors. Digital Watermarking: 9th International Workshop, IWDW 2010, Seoul, Korea, October 1–3, 2010, Revised Selected Papers, volume 6526 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18404-9 (print), 3-642-18405-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18405-5.

Kollias:2006:ANNb

[KSDO06a] Stefanos Kollias, Andreas Stafylopatis, Włodzisław Duch, and Erkki Oja, editors. Artificial Neural Networks — ICANN 2006: 16th International Conference, Athens, Greece, September 10–14, 2006. Proceedings, Part II, volume 4132 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38871-0 (print), 3-540-38873-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
38873-9.

Kollias:2006:ANNa

[KSDO06b] Stefanos D. Kollias, Andreas Stafylopatis, Włodzisław Duch, and Erkki Oja, editors. Artificial Neural Networks — ICANN 2006: 16th International Conference, Athens, Greece, September 10–14, 2006. Proceedings, Part I, volume 4131 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38625-4 (print), 3-540-38627-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38627-8.

Koubarakis:2003:STD

[KSF^{+}03] Manolis Koubarakis, Timos Sellis, Andrew U. Frank, Stéphane Grumbach, Ralf Hartmut Güting, Christian S. Jensen, Nikos Lorentzos, Yannis Manolopoulos, Enrico Nardelli, Barbara Pernici, Hans-Jörg Schek, Michel Scholl, Babis Theodoulidis, and Nectaria Tryfona, editors. Spatio-Temporal Databases: The CHOROCHRONOS Approach, volume 2520 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40552-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 S684 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2520.
htm;   http://www.springerlink.
com/content/978-3-540-
40552-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2520;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b83622.

Kerren:2008:IVH

[KSFN08] Andreas Kerren, John T. Stasko, Jean-Daniel Fekete, and Chris North, editors. Information Visualization: Human-Centered Issues and Perspectives, volume 4950 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70955-X (print), 3-540-70956-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70956-5.

Kuo:2007:EUC

[KSG^{+}07] Tei-Wei Kuo, Edwin Sha, Minyi Guo, Laurence T. Yang, and Zili Shao, editors. Embedded and Ubiquitous Computing: International Conference, EUC 2007, Taipei, Taiwan, December 17–20, 2007. Proceedings, volume 4808 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77091-7 (print), 3-540-77092-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77092-3.

Kadirkamanathan:2009:PRB

[KSG^{+}09] Visakan Kadirkamanathan, Guido Sanguinetti, Mark Girolami, Mahesan Niranjan, and Josselin Noirel, editors. Pattern Recognition in Bioinformatics: 4th IAPR International Conference, PRIB 2009, Sheffield, UK, September 7–9, 2009. Proceedings, volume 5780 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04030-6 (print), 3-642-04031-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04031-3.

Kuhne:2010:MFA

[KSGT10] Thomas Kühne, Bran Selic, Marie-Pierre Gervais, and François Terrier, editors. Modelling Foundations and Applications: 6th European Conference, ECMFA 2010, Paris, France, June 15–18, 2010. Proceedings, volume 6138 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13594-3 (print), 3-642-13595-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13595-8.

Kropatsch:2005:PRD

[KSH05] Walter G. Kropatsch, Robert Sablatnig, and Allan Hanbury, editors. Pattern recognition: 27th DAGM symposium, Vienna, Austria, August 31–September 2, 2005. proceedings, volume 3663 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28703-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28703-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3663.

Krishnamachari:2009:DCS

[KSHM09] Bhaskar Krishnamachari, Subhash Suri, Wendi Heinzelman, and Urbashi Mitra, editors. Distributed Computing in Sensor Systems: 5th IEEE International Conference, DCOSS 2009, Marina del Rey, CA, USA, June 8–10, 2009. Proceedings, volume 5516 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02084-4 (print), 3-642-02085-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02085-8.

Kuznetsov:2011:RSF

[KŚHM11] Sergei O. Kuznetsov, Dominik Ślezak, Daryl H.

Hepting, and Boris G. Mirkin, editors. Rough Sets, Fuzzy Sets, Data Mining and Granular Computing: 13th International Conference, RSFDGrC 2011, Moscow, Russia, June 25–27, 2011. Proceedings, volume 6743 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21880-6 (print), 3-642-21881-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21881-1.

Kahle:2001:PTC

[KSHS01] Reinhard Kahle, Peter Joseph Schroeder-Heister, and Robert F. Stärk, editors. Proof theory in computer science: international seminar, PTCS 2001, Dagstuhl Castle, Germany, October 7–12, 2001: Proceedings, volume 2183 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42752-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 P753 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2183.htm;
 http://www.springerlink.
com/content/978-3-540-
42752-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2183.

Khosrovshahi:2002:TAC

[KSS02] Gholamreza B. Khosrovshahi, Ali Shokoufandeh, and Amin Shokrollahi, editors. Theoretical aspects of computer science: advanced lectures, volume 2292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43328-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .S866 2002; QA76 .T426 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2292.htm;
 http://www.springerlink.
com/content/978-3-540-
43328-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2292.

Papers from the First Summer School on Theoretical Aspects of Computer Science, July 3–10, 2000.

Kapetanios:2008:NLI

[KSS08] Epaminondas Kapetanios, Vijayan Sugumaran, and Myra Spiliopoulou, editors. Natural Language and Information Systems: 13th International Conference on Applications of Natural Language to Information Systems, NLDB 2008 London, UK, June 24–27, 2008 Proceedings, volume 5039 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69857-4 (print), 3-540-69858-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69858-6.

Kuba:1999:IPM

[KSTP99] Attila Kuba, Martin Sámal, and Andrew Todd-Pokropek, editors. Information processing in medical imaging: 16th International Conference, IPMI’99, Visegrad, Hungary, June 28– July 2, 1999: Proceedings, volume 1613 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66167-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN WN 200 I61-2 1999i. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1613.htm;
 http://www.springerlink.
com/content/978-3-540-
66167-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1613.

Kivinen:2011:ALT

[KSUZ11] Jyrki Kivinen, Csaba Szepesvári, Esko Ukkonen, and Thomas Zeugmann, editors. Algorithmic Learning Theory: 22nd International Conference, ALT 2011, Espoo, Finland, October 5–7, 2011. Proceedings, volume 6925 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24411-4 (print), 3-642-24412-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24412-4.

Kreczmar:1990:LRP

[KSW90] Antoni Kreczmar, Andrzej Salwicki, and Marek Warpechowski. LOGLAN ’88–report on the programming language, volume 414 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52325-1 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 133 pp. LCCN QA76.73.L625 K74 1990. DM28.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0414.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        52325-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=414.

Klusch:1999:CIA

[KSW99] Matthias Klusch, Onn M. Shehory, and Gerhard Weiss, editors. Cooperative information agents III: Third International Workshop, CIA’99, Uppsala, Sweden, July 31– August 2, 1999: Proceedings, volume 1652 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66325-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 C53 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1652.htm;
 http://www.springerlink.
com/content/978-3-540-
66325-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1652.

Klein:2003:CSP

[KSW03] Rolf Klein, H.-W. (Hans-Werner) Six, and Lutz Wegner, editors. Computer Science in Perspective: Essays Dedicated to Thomas Ottmann, volume 2598 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00579-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 355 pp. LCCN QA76 .C5476 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2598.htm;
 http://www.springerlink.
com/content/978-3-540-
00579-7;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2598.

Kimmel:2005:SSP

[KSW05] Ron Kimmel, Nir Sochen, and Joachim Weickert, editors. Scale space and PDE methods in computer vision: 5th international conference, Scale-Space 2005, Hofgeismar, Germany, April 7– 9, 2005: Proceedings, volume 3459 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25547-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .S38. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25547-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3459;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b107185.

Kasturi:1996:GRM

[KT96] Rangachar Kasturi and Karl Tombre, editors. Graphics recognition: methods and applications: first International Workshop, University Park, PA, USA, August 10–11, 1995: Selected Papers, volume 1072 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61226-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .G73 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1072.htm;
 http://www.springerlink.
com/content/978-3-540-
61226-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1072.

Kim:2006:MCN

[KT06] Young-Tak Kim and Makoto Takano, editors. Management of Convergence Networks and Services: 9th Asia-Pacific Network Operations and Management Symposium, APNOMS 2006 Busan, Korea, September 27–29, 2006 Proceedings, volume 4238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45776-3 (print), 3-540-46233-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        46233-0.

Krahmer:2010:EMN

[KT10] Emiel Krahmer and Mariët Theune, editors. Empirical Methods in Natural Language Generation: Data-oriented Methods and Empirical Evaluation, volume 5790 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15572-3 (print), 3-642-15573-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15573-4.

Kranzlmuller:2011:ICT

[KT11] Dieter Kranzlmüller and A. Min Toja, editors. Information and Communication on Technology for the Fight against Global Warming: First International Conference, ICT-GLOW 2011, Toulouse, France, August 30–31, 2011. Proceedings, volume 6868 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23446-1 (print), 3-642-23447-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23447-7.

Kolehmainen:2009:ANC

[KTB09] Mikko Kolehmainen, Pekka Toivanen, and Bartlomiej Beliczynski, editors. Adaptive and Natural Computing Algorithms: 9th International Conference, ICANNGA 2009, Kuopio, Finland, April 23–25, 2009, Revised Selected Papers, volume 5495 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04920-6 (print), 3-642-04921-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04921-7.

Keijzer:2005:GPE

[KTC^{+}05] Maarten Keijzer, Andrea Tettamanzi, Pierre Collet, Jano van Hemert, and Marco Tomassini, editors. Genetic programming: 8th European conference, EuroGP 2005 Lausanne, Switzerland, March 30– April 1, 2005 Proceedings, volume 3447 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25436-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.623 .E94 2005. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3447;   http:
//www.springerlink.
com/content/978-3-540-
25436-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3447;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107383.

Karlgren:2009:SPI

[KTH09] Jussi Karlgren, Jorma Tarhio, and Heikki Hyyrö, editors. String Processing and Information Retrieval: 16th International Symposium, SPIRE 2009 Saariselkä, Finland, August 25– 27, 2009 Proceedings, volume 5721 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03783-6 (print), 3-642-03784-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03784-9.

Krommer:1994:NIA

[KU94] Arnold R. Krommer and Christoph W. Ueberhuber. Numerical integration on advanced computer systems, volume 848 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58410-2 (Berlin), 0-387-58410-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 341 pp. LCCN QA299.3 .K76 1994. DM72.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0848.htm;
http://www.springerlink.
com/content/978-3-540-
58410-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=848.

Kuchen:2001:FLP

[KU01] Herbert Kuchen and Kazunori Ueda, editors. Functional and logic programming: 5th International Symposium, FLOPS 2001, Tokyo, Japan, March 7–9, 2001:

Proceedings, volume 2024 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41739-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .F56 2001; QA267.A1 L43 no.2024. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2024.htm;
 http://www.springerlink.
com/content/978-3-540-
41739-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2024.

Kralovic:2006:MFC

[KU06] Rastislav Královič and Paweł Urzyczyn, editors. Mathematical Foundations of Computer Science 2006: 31st International Symposium, MFCS 2006, Stará Lesná, Slovakia, August 28–September 1, 2006. Proceedings, volume 4162 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37791-3 (print), 3-540-37793-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
37793-1.

Kucherov:2009:CPM

[KU09] Gregory Kucherov and Esko Ukkonen, editors. Combinatorial Pattern Matching: 20th Annual Symposium, CPM 2009 Lille, France, June 22– 24, 2009 Proceedings, volume 5577 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02440-8 (print), 3-642-02441-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02441-2.

Koutny:2012:CCT

[KU12] Maciej Koutny and Irek Ulidowski, editors. CON-CUR 2012 — Concurrency Theory: 23rd International Conference, CONCUR 2012, Newcastle upon Tyne, UK, September 4–7, 2012. Proceedings, volume 7454 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32939-X (print), 3-642-32940-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32940-1.

Kano:2000:DCG

[KUA00] Mikio Kano, Masatsugu Urabe, and J. Akiyama, editors. Discrete and Computational Geometry: Second Japanese Conference, JCDCG’98, Tokyo, Japan, December 9–12, 1998: Revised Papers, volume 1763 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67181-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1763.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1763.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67181-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1763.

Kucera:2002:GTC

[Kuc̆02] Ludĕk Kuc̆era, editor. Graph-Theoretic Concepts in Computer Science: 28th International Workshop, WG 2002,: Český Krumlov, Czech Republic, June 13– 15, 2002. Revised Papers, volume 2573 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00331-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G7172 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2573.htm;
 http://www.springerlink.
com/content/978-3-540-
00331-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2573.

Kuchen:2011:FCL

[Kuc11] Herbert Kuchen, editor. Functional and Constraint Logic Programming: 20th International Workshop, WFLP 2011, Odense, Denmark, July 19th, Proceedings, volume 6816 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22530-6 (print), 3-642-22531-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        22531-4.

Kuhne:2007:MSE

[Küh07] Thomas Kühne, editor. Models in Software Engineering: Workshops and Symposia at MoDELS 2006, Genoa, Italy, October 1–6, 2006, Reports and Revised Selected Papers, volume 4364 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69488-9 (print), 3-540-69489-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69489-2.

Kuhlmann:2010:DSL

[Kuh10] Marco Kuhlmann, editor. Dependency Structures and Lexicalized Grammars: An Algebraic Approach, volume 6270 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CO-DEN LNCSD9. ISBN 3-642-14567-1 (print), 3-642-14568-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 125 (est.) pp. LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-642-
14568-1.

Kuich:1992:ALP

[Kui92] Werner Kuich, editor.

Automata, languages, and programming: 19th international colloquium, Wien, Austria, July 13–17, 1992: proceedings, volume 623 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55719-9, 0-387-55719-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.623.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0623.
htm;   http://www.springerlink.
com/content/978-3-540-
55719-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=623.

Proceedings of ICALP 92, the 19th International Colloquium on Automata, Languages, and Programming, sponsored by the European Association for Theoretical Computer Science.

Kullmann:2009:TAS

[Kul09] Oliver Kullmann, editor.

Theory and Applications of Satisfiability Testing

— SAT 2009: 12th International Conference, SAT 2009, Swansea, UK, June 30 — July 3, 2009. Proceedings, volume 5584 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02776-8 (print), 3-642-02777-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02777-2.

Kumar:1990:CTS

[Kum90]  D. Kumar.  Current

trends in SNePS — semantic network processing system, volume 437 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52626-9 (Berlin), 0-387-52626-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). 162 pp. LCCN Q336 .S54 1989. DM33.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0437.
htm;   http://www.springerlink.
com/content/978-3-540-
52626-1;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=437.

Kunii:1984:VEB

[Kun84] Toshiyasu Kunii. VLSI engineering: beyond software engineering, volume 163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CODEN LNCSD9. ISBN 0-387-70002-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 308 pp. LCCN TK7888.4 .V56 1984. US$14.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0163.htm;
http://www.springerlink.
com/content/978-0-387-
70002-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=163.

Kurosawa:2007:ACA

[Kur07] Kaoru Kurosawa, editor. Advances in Cryptology — ASIACRYPT 2007: 13th International Conference on the Theory and Application of Cryptology and Information Security, Kuching, Malaysia, December 2–6, 2007. Proceedings, volume 4833 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76899-8 (print), 3-540-76900-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76900-2.

Kurosu:2009:HCD

[Kur09] Masaaki Kurosu, editor. Human Centered Design: First International Conference, HCD 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009 Proceedings, volume 5619 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02805-5 (print), 3-642-02806-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02806-9.

Kurosawa:2010:ITS

[Kur10] Kaoru Kurosawa, editor. Information Theoretic Security: 4th International Conference, ICITS 2009, Shizuoka, Japan, December 3–6, 2009. Revised Selected Papers, volume 5973 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14495-0 (print), 3-642-14496-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14496-7.

Kurosu:2011:HCD

[Kur11] Masaaki Kurosu, editor. Human Centered Design: Second International Conference, HCD 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011. Proceedings, volume 6776 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21752-4 (print), 3-642-21753-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21753-1.

Kusters:2001:NSI

[Küs01]  Ralf Küsters.  Non-

Standard Inferences in Description Logics, volume 2100 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42397-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 250 pp. LCCN QA267.A1 L43 no.2100; Q334 .K87 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2100.htm;
 http://www.springerlink.
com/content/978-3-540-
42397-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2100.

Kutten:1998:DCI

[Kut98] Shay Kutten, editor. Distributed computing: 12th International Symposium, DISC ’98, Andros, Greece, September 24–26, 1998: Proceedings, volume 1499 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65066-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1499. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1499.htm;
 http://www.springerlink.

com/content/978-3-540-
65066-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1499.

Kutulakos:2012:TTCa

[Kut12a] Kiriakos N. Kutulakos, editor. Trends and Topics in Computer Vision: ECCV 2010 Workshops, Heraklion, Crete, Greece, September 10–11, 2010, Revised Selected Papers, Part I, volume 6553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35748-2 (print), 3-642-35749-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35749-7.

Kutulakos:2012:TTCb

[Kut12b] Kiriakos N. Kutulakos, editor. Trends and Topics in Computer Vision: ECCV 2010 Workshops, Heraklion, Crete, Greece, September 10–11, 2010, Revised Selected Papers, Part II, volume 6554 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35739-3 (print), 3-642-35740-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        35740-4.

Karsai:2004:GPC

[KV04] Gabor Karsai and Eelco Visser, editors. Generative Programming and Component Engineering: Third International Conference, GPCE 2004, Vancouver, Canada, October 24–28, 2004, Proceedings, volume 3286 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23580-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3286.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23580-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3286;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b101929.

Kordon:2008:RST

[KV08] Fabrice Kordon and Tullio Vardanega, editors.

Reliable Software Technologies — Ada-Europe 2008: 13th Ada-Europe International Conference on Reliable Software Technologies, Venice, Italy, June 16–20, 2008. Proceedings, volume 5026 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68621-5 (print), 3-540-68624-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68624-8.

Kulikov:2011:CST

[KV11] Alexander Kulikov and Nikolay Vereshchagin, editors. Computer Science — Theory and Applications: 6th International Computer Science Symposium in Russia, CSR 2011, St. Petersburg, Russia, June 14–18, 2011. Proceedings, volume 6651 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20711-1 (print), 3-642-20712-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20712-9.

Koornneef:2000:CSR

[KvdM00] Floor Koornneef and Meine van der Meulen, editors. Computer Safety, Reliability and Security: 19th International Conference, SAFE-COMP 2000, Rotterdam, The Netherlands, October 24–27, 2000. Proceedings, volume 1943 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41186-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1943. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1943.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41186-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1943.

Kappers:2010:HGPb

[KvETvdH10a] Astrid M. L. Kappers, Jan B. F. van Erp, Wouter M. Bergmann Tiest, and Frans C. T. van der Helm, editors. Haptics: Generating and Perceiving Tangible Sensations: International Conference, EuroHaptics 2010, Amsterdam, July 8–10, 2010. Proceedings, volume 6192 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14074-2 (print), 3-642-14075-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14075-4.

Kappers:2010:HGPa

[KvETvdH10b] Astrid M. L. Kappers, Jan B. F. van Erp, Wouter M. Bergmann Tiest, and Frans C. T. van der Helm, editors. Haptics: Generating and Perceiving Tangible Sensations: International Conference, EuroHaptics 2010, Amsterdam, July 8–10, 2010. Proceedings, Part I, volume 6191 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14063-7 (print), 3-642-14064-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14064-8.

Koch:1998:SMI

[KvG98] Reinhard Koch and Luc van Gool, editors. 3D structure from multiple images of large-scale environments: European Workshop, SMILE ’98, Freiburg, Germany, June 6–7, 1998: Proceedings, volume 1506 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65310-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .S62 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1506.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65310-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1506.

Held in conjunction with the Fifth European Conference on Computer Vision 1998 in Freiburg, Germany.

Kowalczyk:2009:SOC

[KVMH09] Ryszard Kowalczyk, Quoc Bao Vo, Zakaria Maamar, and Michael Huhns, editors. Service-Oriented Computing: Agents, Semantics, and Engineering: AAMAS 2009 International Workshop SO-CASE 2009, Budapest, Hungary, May 11, 2009. Proceedings, volume 5907 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10738-9 (print), 3-642-10739-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10739-9.

Kirchner:1990:ALP

[KW90] H. Kirchner and Wolfgang Wechler, editors. Algebraic and logic programming: second international conference, Nancy, France, October 1–3, 1990: proceedings, volume 463 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-53162-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .A4584 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0463.htm;
 http://www.springerlink.
com/content/978-0-387-
53162-5;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=463.

“Proceedings of the Second International Conference on Algebraic and Logic Programming … organized by CRIN and INRIA-Lorraine”– Foreword.

Kuper:1996:CDA

[KW96] Gabriel Kuper and M. (Mark) Wallace, editors. Constraint databases and applications: ESPRIT WG CONTESSA Workshop, Friedrichshafen, Germany, September 8– 9, 1995: proceedings, volume 1034 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60794-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C64 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1034.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60794-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1034.

Klusch:1998:CIA

[KW98] Matthias Klusch and Gerhard Weiss, editors. Cooperative information agents II: learning, mobility, and electronic commerce for information discovery on the Internet: Second International Workshop, CIA ’98, Paris, France, July 1998: Proceedings, volume 1435 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64676-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1435. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1435.htm;
 http://www.springerlink.
com/content/978-3-540-
64676-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1435.

Kaufmann:2001:DGM

[KW01] Michael Kaufmann and Dorothea Wagner, editors. Drawing Graphs: Methods and Models, volume 2025 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42062-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 312 pp. LCCN T385 .D715 2001; QA267.A1 L43 no.2025. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2025.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42062-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2025.

Kaufmann:2007:GDI

[KW07a] Michael Kaufmann and Dorothea Wagner, editors. Graph Drawing: 14th International Symposium, GD 2006, Karlsruhe, Germany, September 18–20, 2006. Revised Papers, volume 4372 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70903-7 (print), 3-540-70904-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70904-6.

Kobti:2007:AAI

[KW07b] Ziad Kobti and Dan Wu, editors. Advances in Artificial Intelligence: 20th Conference of the Canadian Society for Computational Studies of Intelligence, Canadian AI 2007, Montreal, Canada, May 28–30, 2007. Proceedings, volume 4509 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72664-0 (print), 3-540-72665-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72665-4.

Konev:2007:FCS

[KW07c] Boris Konev and Frank Wolter, editors. Frontiers of Combining Systems: 6th International Symposium, FroCoS 2007 Liverpool, UK, September 10–12, 2007 Proceedings, volume 4720 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74620-X (print), 3-540-74621-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74621-8.

Kopp:2010:GEC

[KW10] Stefan Kopp and Ipke Wachsmuth, editors. Gesture in Embodied Communication and Human-Computer Interaction: 8th International Gesture Workshop, GW 2009, Bielefeld, Germany, February 25–27, 2009, Revised Selected Papers, volume 5934 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12552-2 (print), 3-642-12553-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12553-9.

Kambayashi:2001:DWK

[KWA01] Y. Kambayashi, Werner Winiwarter, and Masatoshi Arikawa, editors. Data warehousing and knowledge discovery: Third International Conference, DaWaK 2001, Munich, Germany, September 5– 7, 2001: Proceedings, volume 2114 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42553-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D37 I58 2001; QA267.A1 L43 no.2114.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2114.
htm;   http://www.springerlink.
com/content/978-3-540-
42553-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2114.

Kambayashi:2002:DWK

[KWA02] Y. Kambayashi, Werner Winiwarter, and Masatoshi Arikawa, editors. Data warehousing and knowledge discovery: 4th International Conference, DaWaK 2002, Aix-en-Provence, France, September 4–6, 2002: Proceedings, volume 2454 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44123-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D37 B396 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2454.htm;
 http://www.springerlink.
com/content/978-3-540-
44123-6;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2454.

Katzenbeisser:2012:TTC

[KWC^{+}12] Stefan Katzenbeisser, Edgar Weippl, L. Jean Camp, Melanie Volkamer, Mike Reiter, and Xinwen Zhang, editors. Trust and Trustworthy Computing: 5th International Conference, TRUST 2012, Vienna, Austria, June 13–15, 2012. Proceedings, volume 7344 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30920-8 (print), 3-642-30921-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30921-2.

King:2006:NIPa

[KWCW06a] Irwin King, Jun Wang, Lai-Wan Chan, and DeLiang Wang, editors. Neural Information Processing: 13th International Conference, ICONIP 2006, Hong Kong, China, October 3–6, 2006. Proceedings, Part I, volume 4232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46479-4 (print), 3-540-46480-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46480-8.

King:2006:NIPb

[KWCW06b] Irwin King, Jun Wang, Lai-Wan Chan, and DeLiang Wang, editors. Neural Information Processing: 13th International Conference, ICONIP 2006, Hong Kong, China, October 3–6, 2006. Proceedings, Part II, volume 4233 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46481-6 (print), 3-540-46482-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46482-2.

King:2006:NIPc

[KWCW06c] Irwin King, Jun Wang, Lai-Wan Chan, and DeLiang Wang, editors. Neural Information Processing: 13th International Conference, ICONIP 2006, Hong Kong, China, October 3–6, 2006. Proceedings, Part III, volume 4234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46484-0 (print), 3-540-46485-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46485-3.

Kuhn:2003:SIT

[KWT03] Werner Kuhn, Michael Worboys, and Sabine Timpf, editors. Spatial Information Theory: Foundations of Geographic Information Science: International Conference, COSIT 2003, Ittingen, Switzerland, September 2003: Proceedings, volume 2825 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20148-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.212 .C68 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2825.htm;
 http://www.springerlink.
com/content/978-3-540-
20148-9;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2825;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13481.

Karl:2004:WSN

[KWW04] Holger Karl, Adam Wolisz, and Andreas Willig, editors. Wireless Sensor Networks: First European Workshop, EWSN 2004, Berlin, Germany, January 19– 21, 2004, Proceedings, volume 2920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20825-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7872.D48 E94 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2920.
htm;   http://www.springerlink.
com/content/978-3-540-
20825-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2920;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94854.

Katz:2007:ACN

[KY07a] Jonathan Katz and Moti Yung, editors. Applied Cryptography and Network Security: 5th International Conference, ACNS 2007, Zhuhai, China, June 5–8, 2007. Proceedings, volume 4521 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72737-X (print), 3-540-72738-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72738-5.

Kleijn:2007:PNO

[KY07b] Jetty Kleijn and Alex Yakovlev, editors. Petri Nets and Other Models of Concurrency — ICATPN 2007: 28th International Conference on Applications and Theory of Petri Nets and Other Models of Concurrency, ICATPN 2007, Siedlce, Poland, June 25–29, 2007. Proceedings, volume 4546 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73093-1 (print), 3-540-73094-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73094-1.

Kosowski:2011:SIC

[KY11] Adrian Kosowski and Masafumi Yamashita, editors. Structural Information and Communication Complexity: 18th International Colloquium, SIROCCO 2011, Gdańsk, Poland, June 26–29, 2011. Proceedings, volume 6796 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22211-0 (print), 3-642-22212-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22212-2.

Kim:2007:ISA

[KYL07] Sehun Kim, Moti Yung, and Hyung-Woo Lee, editors. Information Security Applications: 8th International Workshop, WISA 2007, Jeju Island, Korea, August 27–29, 2007, Revised Selected Papers, volume 4867 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77534-X (print), 3-540-77535-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77535-5.

Kou:2001:ECT

[KYT01] Weidong Kou, Yelena Yesha, and Chung Jen Tan, editors. Electronic commerce technologies: Second International Symposium, ISEC 2001, Hong Kong, China, April 26–28, 2001: Proceedings, volume 2040 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41963-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2040; HF5548.32 .I84 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2040.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41963-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2040.

Komorowski:1997:PDM

[KZ97] J. (Jan) Komorowski and Jan Zytkow, editors. Principles of data mining and knowledge discovery: first European symposium, PKDD ’97, Trondheim, Norway, June 24–27, 1997: proceedings, volume 1263 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63223-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3P16 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1263.htm;
 http://www.springerlink.
com/content/978-3-540-
63223-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1263.

Klusch:2001:CIA

[KZ01] Matthias Klusch and Franco Zambonelli, editors. Cooperative information agents V: 5th International Workshop, CIA 2001, Modena, Italy, September 6–8, 2001: Proceedings, volume 2182 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42545-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q337 .C66 2001; QA267.A1 L43 no.2182. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2182.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42545-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2182.

Klette:2004:CIA

[KŽ04] Reinhard Klette and Jovisa Žunić, editors. Combinatorial Image Analysis: 10th International Workshop, IWCIA 2004, Auckland, New Zealand, December 1–3, 2004, Proceedings, volume 3322 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23942-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3322.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23942-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=

0302-9743&volume=3322;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b103936.

Kumar:2010:EPB

[KZ10a] Ajay Kumar and David Zhang, editors. Ethics and Policy of Biometrics: Third International Conference on Ethics and Policy of Biometrics and International Data Sharing, ICEB 2010, Hong Kong, January 4– 5, 2010. Revised Papers, volume 6005 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12594-8 (print), 3-642-12595-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12595-9.

Kutten:2010:SIC

[KŽ10b] Shay Kutten and Janez Žerovnik, editors. Structural Information and Communication Complexity: 16th International Colloquium, SIROCCO 2009, Piran, Slovenia, May 25–27, 2009, Revised Selected Papers, volume 5869 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11475-X (print), 3-642-11476-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11476-2.

Levi:1994:ALP

[LA94] G. Levi and M. Rodriguez Artalejo, editors. Algebraic and logic programming: 4th international conference, ALP ’94, Madrid, Spain, September 14–16, 1994: proceedings, volume 850 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-58431-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .A45843 1994. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0850.
htm;   http://www.springerlink.
com/content/978-0-387-
58431-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=850.

Proceedings of the 4th International Conference on Algebraic and Logic Programming, held in Madrid, September 14– 16, 1994.

Liu:2005:TAC

[LA05] Zhiming Liu and Keijiro Araki, editors. Theoretical aspects of computing: ICTAC 2004: first international colloquium, Guiyang, China, September 20–24, 2004: Revised selected papers, volume 3407 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25304-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5+. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3407;   http:
//www.springerlink.
com/content/978-3-540-
25304-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3407;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107116.

Liu:2012:ACI

[LABM^{+}12] Jing Liu, Cesare Alippi, Bernadette Bouchon-Meunier, Garrison W. Greenwood, and Hussein A. Abbass, editors. Advances in Computational Intelligence: IEEE World Congress on Computational Intelligence, WCCI 2012, Brisbane, Australia, June 10–15, 2012. Plenary/Invited Lectures, volume 7311 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30686-1 (print), 3-642-30687-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30687-7.

Lacroix:2010:RDS

[Lac10] Zoé Lacroix, editor. Resource Discovery: Second International Workshop, RED 2009, Lyon, France, August 28, 2009. Revised Papers, volume 6162 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14414-4 (print), 3-642-14415-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14415-8.

Lin:2010:EPP

[LAF^{+}10] Hai-Xiang Lin, Michael Alexander, Martti Forsell, Andreas Knüpfer, Radu Prodan, Leonel Sousa, and Achim Streit, editors. Euro-Par 2009 — Parallel Processing Workshops: HPPC, HeteroPar, PROPER, ROIA, UNICORE, VHPC, Delft, The Netherlands, August 25–28, 2009, Revised Selected Papers, volume 6043 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14121-8 (print), 3-642-14122-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14122-5.

Lagergren:2005:CGR

[Lag05] Jens Lagergren, editor.

Comparative Genomics: RECOMB 2004 International Workshop, RCG 2004, Bertinoro, Italy, October 16–19, 2004. Revised selected papers, volume 3388 of Lecture Notes in Bioinformatics and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24455-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH447 .C652 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24455-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3388;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105486.

Laih:2003:ACA

[Lai03] Chi Sung Laih, editor. Advances in Cryptology—ASIACRYPT 2003: 9th International Conference on the Theory and Application of Cryptology and Information Security, Taipei, Taiwan, November 30– December 4, 2003: Proceedings, volume 2894 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20592-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I555 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2894.
htm;   http://www.springerlink.
com/content/978-3-540-

20592-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2894;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94617.

Lamb:1996:SSD

[Lam96] David Alex Lamb, editor. Studies of software design: ICSE ’93 workshop, Baltimore, Maryland, USA, May 17– 18, 1993: Selected Papers, volume 1078 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61285-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I33 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1078.
htm;   http://www.springerlink.
com/content/978-3-540-
61285-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1078.

Lambrix:2000:PWR

[Lam00] Patrick Lambrix. Part-Whole Reasoning in an Object-Centered Framework, volume 1771 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67225-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 194 pp. LCCN Q387 .L35 2000. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1771.
htm;   http://www.springerlink.
com/content/978-3-540-
67225-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1771.

Lauer:1993:FPC

[Lau93a] Peter E. Lauer, editor. Functional programming, concurrency, simulation and automated reasoning: international lecture series 1991– 1992, McMaster University, Hamilton, Ontario, Canada, volume 693 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56883-2 (Berlin), 0-387-56883-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.61 .F858 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0693.htm;
 http://www.springerlink.

com/content/978-3-540-
56883-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=693.

Laugier:1993:GRP

[Lau93b] Christian Laugier, editor. Geometric reasoning for perception and action: workshop, Grenoble, France, September 16–17, 1991: Selected Papers, volume 708 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57132-9 (Berlin), 0-387-57132-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G46 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0708.
htm;   http://www.springerlink.
com/content/978-3-540-
57132-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=708.

Laurini:2000:AVI

[Lau00] Robert Laurini, editor.

Advances in Visual Information Systems: 4th International Conference, VISUAL 2000, Lyon, France, November 2–4, 2000. Proceedings, volume 1929 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41177-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.I6 V56 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1929.
htm;   http://www.springerlink.
com/content/978-3-540-
41177-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1929.

Lau:2001:LBP

[Lau01] K.-K. Lau, editor. Logic based program synthesis and transformation: 10th International Workshop, LOPSTER 2000, London, UK, July 24– 28, 2000: Selected Papers, volume 2042 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42127-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I57 2000; QA267.A1 L43 no.2042. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2042.htm;
 http://www.springerlink.
com/content/978-3-540-
42127-6;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2042.

Laud:2012:IST

[Lau12] Peeter Laud, editor. Information Security Technology for Applications: 16th Nordic Conference on Secure IT Systems, NordSec 2011, Tallinn, Estonia, October 26–28, 2011, Revised Selected Papers, volume 7161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29614-9 (print), 3-642-29615-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29615-4.

Lockemann:1978:ISM

[LB78] P. C. Lockemann and G. Bracchi, editors. Information systems methodology: proceedings, 2nd conference of the European Cooperation in Informatics, Venice, October 10–12, 1978, volume 65 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1978. CODEN LNCSD9. ISBN 0-387-08934-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T58.6 .E9 1978. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0065.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        08934-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=65.

Leeser:1990:HSV

[LB90] M. Leeser and G. Brown. Hardware specification, verification, and synthesis: Mathematical aspects: proceedings, volume 408 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-97226-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 402 pp. LCCN TK7874 .H385 1990. DM57.50. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0408.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        97226-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=408.

Lee:2003:IAM

[LB03] Jaeho Lee and Mike Barley, editors. Intelligent Agents and Multi-Agent Systems: 6th Pacific Rim International Workshop on Multi-Agents, PRIMA 2003, Seoul, Korea, November 7–8, 2003: Proceedings, volume 2891 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20460-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2891.
htm;   http://www.springerlink.
com/content/978-3-540-
20460-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2891;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94219.

Lau:2005:FMS

[LB05] K.-K. (Kung-Kiu) Lau and Richard Banach, editors. Formal methods and software engineering: 7th International Conference on Formal Engineering Methods, ICFEM 2005, Manchester, UK, November 1–4, 2005: proceedings, volume 3785 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29797-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 I5788 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        29797-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3785.

Lengauer:2004:DSP

[LBCO04] Christian Lengauer, Don Batory, Charles Consel, and Martin Odersky, editors. Domainspecific program generation: international seminar, Dagstuhl Castle, Germany, March 23– 28, 2003, revised papers, volume 3016 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22119-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        22119-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3016.

Liu:2009:WAS

[LBDW09] Benyuan Liu, Azer Bestavros, Ding-Zhu Du, and Jie Wang, editors. Wireless Algorithms, Systems, and Applications: 4th International Conference, WASA 2009, Boston, MA, USA, August 16–18, 2009. Proceedings, volume 5682 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03416-0 (print), 3-642-03417-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03417-6.

Lee:2010:AFT

[LBM^{+}10] Yeunsook Lee, Z. Zenn Bien, Mounir Mokhtari, Jeong Tai Kim, Mignon Park, Jongbae Kim, Heyoung Lee, and Ismail Khalil, editors. Aging Friendly Technology for Health and Independence: 8th International Conference on Smart Homes and Health Telematics, ICOST 2010, Seoul, Korea, June 22– 24, 2010. Proceedings, volume 6159 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13777-6 (print), 3-642-13778-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13778-5.

Laber:2008:LTI

[LBNF08] Eduardo Sany Laber, Claudson Bornstein, Loana Tito Nogueira, and Luerbio Faria, editors. LATIN 2008: Theoretical Informatics: 8th Latin American Symposium, Búzios, Brazil, April 7–11, 2008. Proceedings, volume 4957 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78772-0 (print), 3-540-78773-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78773-0.

Li:2005:USP

[LBO05] Mingshu Li, Barry Boehm, and Leon J. Osterweil, editors. Unifying the software process spectrum: International Software Process Workshop, SPW 2005, Beijing, China, May 25– 27, 2005: revised selected papers, volume 3840 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-31112-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 I567 2005. URL http:

//www.springerlink.
com/content/978-3-540-
31112-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3840.

Lee:2000:BMC

[LBP00] S.-W. Lee, Heinrich H.

Bülthoff, and Tomaso Poggio, editors. Biologically motivated computer vision: First IEEE/CS International Workshop BMCV 2000, Seoul, Korea, May 2000: Proceedings, volume 1811 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67560-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1811; QP383 .I34 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1811.htm;
 http://www.springerlink.
com/content/978-3-540-
67560-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1811.

Leonardis:2006:CVEa

[LBP06a] Aleš Leonardis, Horst Bischof, and Axel Pinz, editors. Computer Vision — ECCV 2006: 9th European Conference on Computer Vision, Graz, Austria, May 7–13, 2006. Proceedings, Part I, volume 3951 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33832-2 (print), 3-540-33833-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33833-8.

Leonardis:2006:CVEb

[LBP06b] Aleš Leonardis, Horst Bischof, and Axel Pinz, editors. Computer Vision — ECCV 2006: 9th European Conference on Computer Vision, Graz, Austria, May 7–13, 2006. Proceedings, Part II, volume 3952 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33834-9 (print), 3-540-33835-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33835-2.

Leonardis:2006:CVEc

[LBP06c] Aleš Leonardis, Horst Bischof, and Axel Pinz, editors. Computer Vision — ECCV 2006: 9th European Conference on Computer Vision, Graz, Austria, May 7–13, 2006, Proceedings, Part III, volume 3953 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33836-5 (print), 3-540-33837-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33837-6.

Leonardis:2006:CVEd

[LBP06d] Aleš Leonardis, Horst Bischof, and Axel Pinz, editors. Computer Vision — ECCV 2006: 9th European Conference on Computer Vision, Graz, Austria, May 7–13, 2006, Proceedings, Part IV, volume 3954 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33838-1 (print), 3-540-33839-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33839-0.

Lenz:1998:CBR

[LBSBW98] Mario Lenz, Brigitte Bartsch-Spörl, Hans-Dieter Burkhard, and Stefan Wess, editors. Case-Based Reasoning Technology: From Foundations to Applications, volume 1400 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64572-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 405 pp. LCCN QA267.A1 L43 no.1400.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1400.
htm;   http://www.springerlink.
com/content/978-3-540-
64572-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1400.

Contents: Extending some concepts of CBR: foundations of case retrieval nets / Hans-Dieter Burkhard – Diagnosis and decision support / Mario Lenz, Eric Auriol, Michel Manago — Intelligent sales support with CBR / Wolfgang Wilke, Mario Lenz, Stefan Wess – Textual CBR / Mario Lenz, Andre Hubner, Mirjam Kunze – Using configuration techniques for adaptation / Wolfgang Wilke, Barry Smyth, Padraig Cunningham — CBR applied to planning / Ralph Bergmann … [et al.] – CBR for design / Katy Borner — CBR for experimental software engineering / Klaus-Dieter Althoff … [et al.] — CBR for tutoring and help systems / Gerhard Weber, Thomas J. Schult — CBR in medicine / Lothar Gierl, Mathias Bull, Rainer Schmidt — Methodology for building CBR applications / Ralph Bergmann and Klaus-Dieter Althoff — Related areas / Gerd Kamp, Steffen Lange, Christoph Globig.

Lai:2006:ACA

[LC06] Xuejia Lai and Kefei Chen, editors. Advances in Cryptology — ASI-ACRYPT 2006: 12th International Conference on the Theory and Application of Cryptology and Information Security, Shanghai, China, December 3–7, 2006. Proceedings, volume 4284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49475-8 (print), 3-540-49476-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49476-8.

Lee:2009:ISC

[LC09] Pil Joong Lee and Jung Hee Cheon, editors. Information Security and Cryptology — ICISC 2008: 11th International Conference, Seoul, Korea, December 3–5, 2008, Revised Selected Papers, volume 5461 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00729-5 (print), 3-642-00730-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00730-9.

Liu:2011:ICA

[LC11] Baoxiang Liu and Chunlai Chai, editors. Information Computing and Applications: Second International Conference, ICICA 2011, Qinhuangdao, China, October 28– 31, 2011. Proceedings, volume 7030 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25254-0 (print), 3-642-25255-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25255-6.

Liu:1997:AID

[LCB97] X. (Xiaohui) Liu, P. (Paul)

Cohen, and M. (Michael) Berthold, editors. Advances in intelligent data analysis: reasoning about data: Second International Symposium, IDA-97, London, UK, August 1–4, 1997: proceedings, volume 1280 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63346-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA276.A1I28 1997.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t1280.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63346-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1280.

Li:2012:AGP

[LCB12] Ruixuan Li, Jiannong Cao, and Julien Bourgeois, editors. Advances in Grid and Pervasive Computing: 7th International Conference, GPC 2012, Hong Kong, China, May 11–13, 2012. Proceedings, volume 7296 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30766-3 (print), 3-642-30767-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30767-6.

Laender:2009:CMI

[LCD^{+}09] Alberto H. F. Laender, Silvana Castano, Umeshwar Dayal, Fabio Casati, and José Palazzo M. de Oliveira, editors. Conceptual Modeling — ER 2009: 28th International Conference on Conceptual Modeling, Gramado, Brazil, November 9– 12, 2009. Proceedings, volume 5829 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04839-0 (print), 3-642-04840-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04840-1.

Lytras:2008:ETI

[LCDT08] Miltiadis D. Lytras, John M. Carroll, Ernesto Damiani, and Robert D. Tennyson, editors. Emerging Technologies and Information Systems for the Knowledge Society: First World Summit on the Knowledge Society, WSKS 2008, Athens, Greece, September 24– 26, 2008. Proceedings, volume 5288 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87780-0 (print), 3-540-87781-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87781-3.

Liardet:2004:AEI

[LCF^{+}04] Pierre Liardet, Pierre Collet, Cyril Fonlupt, Evelyne Lutton, and Marc Schoenauer, editors. Artificial Evolution: 6th International Conference, Evolution Artificielle, EA 2003, Marseilles, France, October 27–30, 2003, Revised Selected Papers, volume 2936 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21523-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2936.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21523-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2936;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96080.

Luk:1997:FPL

[LCG97] Wayne Luk, Peter Y. K. Cheung, and Manfred Glesner, editors. Fieldprogrammable logic and applications: 7th International Workshop, FPL’97, London, UK, September 1–3, 1997: proceedings, volume 1304 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63465-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.G36 I48 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1304.
htm;   http://www.springerlink.
com/content/978-3-540-
63465-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1304.

Leavens:2012:TAO

[LCH^{+}12] Gary T. Leavens, Shigeru Chiba, Michael Haupt, Klaus Ostermann, and Eric Wohlstadter, editors. Transactions on Aspect-Oriented Software Development IX, volume 7271 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35550-1 (print), 3-642-35551-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35551-6.

Liddell:1996:HPC

[LCHS96] Heather Mary Liddell, A. Colbrook, B. Hertzberger, and P. Sloot, editors. High-performance computing and networking: international conference and exhibition, HPCN EUROPE 1966, Brussels, Belgium, April 15– 19, 1996: proceedings, volume 1067 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61142-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H52 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1067.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61142-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1067.

Lee:2008:CHI

[LCHS08] Seongil Lee, Hyunseung Choo, Sungdo Ha, and In Chul Shin, editors. Computer-Human Interaction: 8th Asia-Pacific Conference, APCHI 2008 Seoul, Korea, July 6–9, 2008 Proceedings, volume 5068 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70584-8 (print), 3-540-70585-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70585-7.

Lang:1991:MSK

[LCS91] Ewald Lang, Kai-Uwe Carstensen, and Geoffrey Simmons. Modelling spatial knowledge on a linguistic basis: theory, prototype, integration, volume 481 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-53718-X (Berlin), 0-387-53718-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 137 pp. LCCN Q387 .L36 1991. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0481.
htm;   http://www.springerlink.
com/content/978-3-540-
53718-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=481.

Liu:2003:IDE

[LCY03] Jiming Liu, Yiuming Cheung, and Hujun Yin, editors. Intelligent Data Engineering and Automated Learning: 4th International Conference, IDEAL 2003 Hong Kong, China, March 21– 23, 2003 Revised Papers, volume 2690 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40550-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I138 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2690.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40550-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2690;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11717.

Lee:2010:FAI

[LCY10] Der-Tsai Lee, Danny Z. Chen, and Shi Ying, editors. Frontiers in Algorithmics: 4th International Workshop, FAW 2010, Wuhan, China, August 11–13, 2010. Proceedings, volume 6213 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14552-3 (print), 3-642-14553-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14553-7.

Luo:2011:NHW

[LCY^{+}11] Xiangfeng Luo, Yiwei Cao, Bo Yang, Jianxun Liu, and Feiyue Ye, editors. New Horizons in Web-Based Learning — ICWL 2010 Workshops: ICWL 2010 Workshops: STEG, CICW, WGLBWS, and IWKDEWL, Shanghai, China, December 7– 11, 2010 Revised Selected Papers, volume 6537 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20538-0 (print), 3-642-20539-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20539-2.

Lavrac:1997:ILP

[LD97] Nada Lavrac and Saso Dzeroski, editors. Inductive logic programming: 7th international workshop, ILP-97, Prague, Czech Republic, September 17–20, 1997: proceedings, volume 1297 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63514-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63.I52 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1297.
htm;   http://www.springerlink.
com/content/978-3-540-
63514-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1297.

Lorenz:2005:NIIb

[LD05a] Pascal Lorenz and Petre Dini, editors. Networking: ICN 2005: 4th international conference on networking, Réunion Island, April 17–21, 2005, proceedings, Part II, volume 3421 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25338-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25338-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3421;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107118.

Lorenz:2005:NIIa

[LD05b] Pascal Lorenz and Petre Dini, editors. Networking: ICN 2005: 4th international conference on networking, Réunion Island, France, April 17– 21, 2005, proceedings, Part I, volume 3420 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25339-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25339-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3420;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107117.

Lytras:2009:VEK

[LDC^{+}09] Miltiadis D. Lytras, Ernesto Damiani, John M.

Carroll, Robert D. Tennyson, David Avison, Ambjörn Naeve, Adrian Dale, Paul Lefrere, Felix Tan, Janice Sipior, and Gottfried Vossen, editors. Visioning and Engineering the Knowledge Society. A Web Science Perspective: Second World Summit on the Knowledge Society, WSKS 2009, Chania, Crete, Greece, September 16–18, 2009. Proceedings, volume 5736 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04753-X (print), 3-642-04754-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04754-1.

Leivant:2007:LLI

[LdQ07] Daniel Leivant and Ruy de Queiroz, editors. Logic, Language, Information and Computation: 14th International Workshop, WoL-LIC 2007, Rio de Janeiro, Brazil, July 2–5, 2007. Proceedings, volume 4576 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73443-0 (print), 3-540-73445-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73445-1.

Langmaack:1994:FTR

[LdRV94] Hans Langmaack, W.-P.

de Roever, and J. Vytopil, editors. Formal techniques in real-time and fault-tolerant systems: Third International Symposium organized jointly with the Working Group Provably Correct Systems, Pro-CoS, Lubeck, Germany, September 19–23, 1994: proceedings, volume 863 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-58468-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54 .F66 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0863.htm;
 http://www.springerlink.
com/content/978-0-387-
58468-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=863.

Lindemann:2004:MST

[LDTU04]  Gabriela  Lindemann,

Jörg Denzinger, Ingo J. Timm, and Rainer Unland, editors. Multiagent system technologies: second German conference, MATES 2004, Erfurt, Germany, September 29–30, 2004. Proceedings, volume 3187 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23222-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        23222-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3187.

Liu:2010:IRAa

[LDXZ10a] Honghai Liu, Han Ding, Zhenhua Xiong, and Xiangyang Zhu, editors. Intelligent Robotics and Applications: Third International Conference, ICIRA 2010, Shanghai, China, November 10–12, 2010. Proceedings, Part I, volume 6424 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16583-4 (print), 3-642-16584-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16584-9.

Liu:2010:IRAb

[LDXZ10b] Honghai Liu, Han Ding, Zhenhua Xiong, and Xiangyang Zhu, editors. Intelligent Robotics and Applications: Third International Conference, ICIRA 2010, Shanghai, China, November 10–12, 2010. Proceedings, Part II, volume 6425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16586-9 (print), 3-642-16587-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16587-0.

LeCharlier:1994:SAF

[Le 94] Baudouin Le Charlier, editor. Static analysis: First International Static Analysis Symposium, SAS ’94, Namur, Belgium, September 28–30, 1994: proceedings, volume 864 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58485-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0864.htm;
 http://www.springerlink.
com/content/978-3-540-
58485-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=864.

Spine title: Static analysis symposium.

LeMetayer:2002:PLS

[Le 02] D. Le Métayer, editor. Programming languages and systems: 11th European Symposium on Programming, ESOP 2002, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2002, Grenoble, France, April 8–12, 2002: Proceedings, volume 2305 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43363-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2305.htm;
http://www.springerlink.
com/content/978-3-540-
43363-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2305.

Lee:2001:GCG

[Lee01] C. A. Lee, editor. Grid computing — GRID 2001: Second International Workshop, Denver, CO, USA, November 12, 2001: Proceedings, volume 2242 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42949-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 G74 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2242.
htm;   http://www.springerlink.
com/content/978-3-540-
42949-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2242.

Lee:2004:ACA

[Lee04] Pil Joong Lee, editor.

Advances in cryptology, ASIACRYPT 2004: 10th International Conference on the Theory and Application of Cryptology and Information Security, Jeju Island, Korea, December 5–9, 2004: Proceedings, volume 3329 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23975-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I555 2004; QA75.5 .L48 no. 3329. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3329;   http:
//www.springerlink.
com/content/978-3-540-
23975-8;   http://www.
springerlink.com/openurl.
asp?genre=volume&id=
doi:10.1007/b104116.

Lee:2011:PPC

[Lee11] Jimmy Lee, editor. Principles and Practice of Constraint Programming

– CP 2011: 17th International Conference, CP 2011, Perugia, Italy, September 12–16, 2011. Proceedings, volume 6876 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23785-1 (print), 3-642-23786-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23786-7.

Lehnert:2011:EAC

[Leh11] Ralf Lehnert, editor.

Energy-Aware Communications: 17th International Workshop, EU-NICE 2011, Dresden, Germany, September 5– 7, 2011. Proceedings, volume 6955 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23540-9 (print), 3-642-23541-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23541-2.

Leilich:1974:GNF

[Lei74] Hans-Otto Leilich, editor. GI-NTG Fachtagung Struktur und Betrieb von Rechensystemen, Braunschweig, 20.-22.3.1974. (German) [GI-NTG Technical Conference on the Structure and Business of Computer Systems], volume 8 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CO-DEN LNCSD9. ISBN 3-540-06677-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.8.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0008.
htm;   http://www.springerlink.
com/content/978-3-540-
06677-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=8. En-

glish or German.

Leivant:1995:LCC

[Lei95] Daniel Maurice Raphael Leivant, editor. Logic and computational complexity: international workshop, LCC ’94, Indianapolis, IN, USA, October 13–16, 1994: Selected Papers, volume 960 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60178-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.A1 L625 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0960.htm;
 http://www.springerlink.
com/content/978-3-540-
60178-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=960.

Leman:1997:MGC

[Lem97] Marc Leman, editor. Music, gestalt, and computing: studies in cognitive and systematic musicology, volume 1317 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63526-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ML3830 .M975 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1317.htm;
http://www.springerlink.
com/content/978-3-540-
63526-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1317.

Papers from the Joint International Conference on Cognitive and Systematic Musicology, held Sept. 8–11, 1996, at the College of Europe, Brugge, Belgium.

Lenz:1990:GTM

[Len90] Reiner Lenz. Group theoretical methods in image processing, volume 413 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52290-5 (USA) (Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 138 pp. LCCN TA1632 .L46 1990. DM28.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0413.
htm;   http://www.springerlink.
com/content/978-0-387-
52290-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=413.

“Based on a series of lectures [given] at the Linköping University Department of Electrical Engineering in 1988”– Foreword.

Lengauer:1993:AEF

[Len93] T. Lengauer, editor. Algorithms — ESA ’93: first annual European symposium, Bad Honnef, Germany, September 30– October 2, 1993: proceedings, volume 726 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57273-2 (Berlin), 0-387-57273-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0726.htm;

http://www.springerlink.
com/content/978-3-540-
57273-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=726.

Lenzerini:1997:AIA

[Len97] Maurizio Lenzerini, editor. AI IA 97, advances in artificial intelligence: 5th Congress of the Italian Association for Artificial Intelligence, Rome, Italy, September 17–19, 1997: proceedings, volume 1321 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63576-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A88 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1321.htm;
http://www.springerlink.
com/content/978-3-540-
63576-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1321.

Lenski:2004:LVA

[Len04] Wolfgang Lenski, editor.

Logic versus Approximation: Essays Dedicated to Michael M. Richter on the Occasion of his 65th Birthday, volume 3075 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22562-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3075.htm;
 http://www.springerlink.
com/content/978-3-540-
22562-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3075;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98383.

Leong:2000:MDA

[Leo00] Hong Va Leong, editor. Mobile data access: First International Conference, MDA ’99, Hong Kong, China, December 16–17, 1999: Proceedings, volume 1748 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-66878-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1748. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t1748.htm;
http://www.springerlink.
com/content/978-3-540-
66878-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1748.

Leonardi:2004:AMW

[Leo04] Stefano Leonardi, editor. Algorithms and Models for the Web-Graph: Third International Workshop, WAW 2004, Rome, Italy, October 16, 2004, Proceedings, volume 3243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23427-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W63 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3243.
htm;   http://www.springerlink.
com/content/978-3-540-
23427-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3243;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101552.

Leonardi:2009:INE

[Leo09] Stefano Leonardi, editor.

Internet and Network Economics: 5th International Workshop, WINE 2009, Rome, Italy, December 14–18, 2009. Proceedings, volume 5929 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10840-7 (print), 3-642-10841-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10841-9.

Liao:2010:MIA

[LEP^{+}10] Hongen Liao, P. J. ‘Eddie’ Edwards, Xiaochuan Pan, Yong Fan, and Guang-Zhong Yang, editors. Medical Imaging and Augmented Reality: 5th International Workshop, MIAR 2010, Beijing, China, September 19–20, 2010. Proceedings, volume 6326 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15698-3 (print), 3-642-15699-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15699-1.

Lescanne:1987:RTA

[Les87] Pierre Lescanne, editor. Rewriting techniques and applications: Bordeaux, France, May 25–27, 1987, proceedings, volume 256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17220-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .R481 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0256.htm;
http://www.springerlink.
com/content/978-0-387-
17220-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=256.

Proceedings of the Second International Conference on Rewriting Techniques and Applications.

Leung:1997:VIS

[Leu97] Clement H. C. Leung, editor. Visual information systems, volume 1306 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63636-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .V58 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1306.htm;
 http://www.springerlink.
com/content/978-3-540-
63636-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1306.

Based on the First International Conference on Visual Information Systems, held in February 1996.

Leuschel:2003:LBP

[Leu03] Michael Leuschel, editor. Logic based program synthesis and transportation: 12th International Workshop, LOPSTR 2002, Madrid, Spain, September 17–20, 2002: revised Selected Papers, volume 2664 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40438-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I57 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2664.htm;
 http://www.springerlink.
com/content/978-3-540-
40438-5;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2664.

Leucker:2008:RVI

[Leu08] Martin Leucker, editor. Runtime Verification: 8th International Workshop, RV 2008, Budapest, Hungary, March 30, 2008. Selected Papers, volume 5289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89246-X (print), 3-540-89247-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89247-2.

Levene:1992:NUR

[Lev92] M. (Mark) Levene. The nested universal relation database model, volume 595 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55493-9 (Berlin), 0-387-55493-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). 177 pp. LCCN QA76.9.D3 L48 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0595.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55493-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=595.

Revision of the author’s thesis (Ph. D.) — Birkbeck College, 1990.

Levi:1998:SAI

[Lev98] G. (Giorgio) Levi, editor. Static analysis: 5th International Symposium, SAS’98, Pisa, Italy, September 14–16, 1998: Proceedings, volume 1503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65014-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6.I5733 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1503.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65014-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1503.

Leyton:2001:GTS

[Ley01] Michael Leyton. A generative theory of shape: questioning Klein’s Erlanger program, volume 2145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42717-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 554 pp. LCCN T57.825 .L49 2001; QA267.A1 L43 no.2145. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2145.htm;
 http://www.springerlink.
com/content/978-3-540-
42717-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2145.

Also available via the World Wide Web with additional title: A Generative Theory of Shape.

Liu:2007:ANNa

[LFH^{+}07a] Derong Liu, Shumin Fei, Zeng-Guang Hou, Huaguang Zhang, and Changyin Sun, editors. Advances in Neural Networks — ISNN 2007: 4th International Symposium on Neural Networks, ISNN 2007, Nanjing, China, June 3–7, 2007, Proceedings, Part I, volume 4491 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72382-X (print), 3-540-72383-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72383-7.

Liu:2007:ANNb

[LFH^{+}07b] Derong Liu, Shumin Fei, Zengguang Hou, Huaguang Zhang, and Changyin Sun, editors. Advances in Neural Networks — ISNN 2007: 4th International Symposium on Neural Networks, ISNN 2007, Nanjing, China, June 3–7, 2007, Proceedings, Part II, volume 4492 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72392-7 (print), 3-540-72393-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72393-6.

Liu:2007:ANNc

[LFH^{+}07c] Derong Liu, Shumin Fei, Zengguang Hou, Huaguang Zhang, and Changyin Sun, editors. Advances in Neural Networks — ISNN 2007: 4th International Symposium on Neural Networks, ISNN 2007, Nanjing, China, June 3–7, 2007, Proceedings, Part III, volume 4493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72394-3 (print), 3-540-72395-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72395-0.

Li:2007:BIC

[LFIM07] Kang Li, Minrui Fei, George William Irwin, and Shiwei Ma, editors. Bio-Inspired Computational Intelligence and Applications: International Conference on Life System Modeling and Simulation, LSMS 2007, Shanghai, China, September 14–17, 2007. Proceedings, volume 4688 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74768-0 (print), 3-540-74769-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74769-7.

Li:2010:LSMa

[LFJI10a] Kang Li, Minrui Fei, Li Jia, and George W. Irwin, editors. Life System Modeling and Intelligent Computing: International Conference on Life System Modeling and Simulation, LSMS 2010, and International Conference on Intelligent Computing for Sustainable Energy and Environment, ICSEE 2010, Wuxi, China, September 17–20, 2010, Proceedings, Part I, volume 6328 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15620-7 (print), 3-642-15621-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15621-2.

Li:2010:LSMb

[LFJI10b] Kang Li, Minrui Fei, Li Jia, and George W. Irwin, editors. Life System Modeling and Intelligent Computing: International Conference on Life System Modeling and Simulation, LSMS 2010, and International Conference on Intelligent Computing for Sustainable Energy and Environment, IC-SEE 2010, Wuxi, China, September 17–20, 2010, Proceedings, Part II, volume 6329 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15596-0 (print), 3-642-15597-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15597-0.

Lim:2002:DLP

[LFK^{+}02] Ee-Peng Lim, Schubert Foo, Chris Khoo, Hsinchun Chen, Edward Fox, Shalini Urs, and Thanos Costantino, editors. Digital Libraries: People, Knowledge, and Technology: 5th International Conference on Asian Digital Libraries, ICADL 2002, Singapore, December 11–14, 2002. Proceedings, volume 2555 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00261-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4080 .I578 2002. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2555.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00261-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2555.

Also available via the

World Wide Web.

Losada:2005:AIR

[LFL05] David E. Losada and Juan M. Fernández-Luna, editors. Advances in information retrieval: 27th European conference on IR research, ECIR 2005, Santiago de Compostela, Spain, March 21–23, 2005, proceedings, volume 3408 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25295-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 E92 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25295-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3408;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b107096.

Li:2009:ADW

[LFP^{+}09] Qing Li, Ling Feng, Jian Pei, Sean X. Wang, Xiaofang Zhou, and Qiao-Ming Zhu, editors. Advances in Data and Web Management: Joint International Conferences, APWeb/WAIM 2009 Suzhou, China, April 2–4, 2009 Proceedings, volume 5446 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00671-X (print), 3-642-00672-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00672-2.

Lee:1994:DID

[LG94] John P. Lee and Georges G.

Grinstein, editors. Database issues for data visualization: IEEE Visualization ’93 workshop, San Jose, California, USA, October 26, 1993: proceedings, volume 871 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58519-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D35853 1994.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0871.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        58519-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=871.

Little:1995:NOS

[LG95] Thomas Little and Riccardo Gusella, editors. Network and operating system support for digital audio and video: 5th International Workshop, NOSSDAV ’95, Durham, New Hampshire, USA, April 19–21, 1995: proceedings, volume 1018 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60647-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7881.4 .N68 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1018.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60647-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1018.

Leopold:1999:MAS

[LG99] Helmut Leopold and Narciso Garcia, editors. Multimedia applications, services and techniques, ECMAST ’99: 4th European Conference, Madrid, Spain, May 26–28, 1999: Proceedings, volume 1629 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66082-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .E25 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1629.htm;
 http://www.springerlink.
com/content/978-3-540-
66082-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1629.

Lowe:2005:WEI

[LG05] David Lowe and Martin Gaedke, editors. Web engineering: 5th international conference, ICWE 2005, Sydney, Australia, July 27–29, 2005: proceedings, volume 3579 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27996-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .I28 2005eb; TK5105.888; TK5105.888 .I28 2005; Internet. URL

                                        http://www.springerlink.
                                        com/content/978-3-540-
                                        27996-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3579.

Lavrac:2003:MLE

[LGBT03] Nada Lavrač, Dragan Gamberger, Hendrik Blockeel, and Ljupčo Todorovski, editors. Machine Learning: ECML 2003: 14th European Conference on Machine Learning, Cavtat-Dubrovnik, Croatia, September 22– 26, 2003: Proceedings, volume 2837 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20121-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2837.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20121-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2837;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13633.

Lengauer:1997:EPP

[LGG97] Christian Lengauer, Martin Griebl, and Sergei Gorlatch, editors. Euro-Par’97, parallel processing: Third International Euro-Par Conference, Passau, Germany, August 26–29, 1997: proceedings, volume 1300 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63440-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I5535 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1300.htm;
http://www.springerlink.
com/content/978-3-540-
63440-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1300.

Lagana:2004:CSAb

[LGK^{+}04a] Antonio Laganà, Marina L. Gavrilova, Vipin Kumar, Youngsong Mun, C. J. Kenneth Tan, and Osvaldo Gervasi, editors. Computational Science and Its Applications

— ICCSA 2004: International Conference, Assisi, Italy, May 14– 17, 2004, Proceedings, Part II, volume 3044 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22056-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3044.htm;
 http://www.springerlink.
com/content/978-3-540-
22056-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3044;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98051.

Lagana:2004:CSAc

[LGK^{+}04b] Antonio Laganà, Marina L. Gavrilova, Vipin Kumar, Youngsong Mun, C. J. Kenneth Tan, and Osvaldo Gervasi, editors. Computational Science and Its Applications

— ICCSA 2004: International Conference, Assisi, Italy, May 14– 17, 2004, Proceedings, Part III, volume 3045 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22057-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3045.htm;
 http://www.springerlink.
com/content/978-3-540-
22057-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3045;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98053.

Lagana:2004:CSAa

[LGK^{+}04c] Antonio Laganà, Marina L. Gavrilova, Vipin Kumar, Youngsong Mun, C. J. Kenneth Tan, and Osvaldo Gervasi, editors. Computational Science and Its Applications— ICCSA 2004: International Conference, Assisi, Italy, May 14– 17, 2004: Proceedings, volume 3043 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22054-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 P73 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3043.
htm;   http://www.springerlink.
com/content/978-3-540-
22054-1;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3043;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98048.

Lagana:2004:CSAd

[LGK^{+}04d] Antonio Laganà, Marina L. Gavrilova, Vipin Kumar, Youngsong Mun, C. J. Kenneth Tan, and Osvaldo Gervasi, editors. Computational Science and Its Applications - ICCSA 2004: International Conference, Assisi, Italy, May 14– 17, 2004, Proceedings, Part IV, volume 3046 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22060-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3046.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22060-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3046;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98054.

Lozano:2011:AAI

[LGM11] Jose A. Lozano, José A.

Gámez, and José A. Moreno, editors. Advances in Artificial Intelligence: 14th Conference of the Spanish Association for Artificial Intelligence, CAEPIA 2011, La Laguna, Spain, November 7–11, 2011. Proceedings, volume 7023 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25273-7 (print), 3-642-25274-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25274-7.

Lucena:2004:SEM

[LGR^{+}04] Carlos Lucena, Alessandro Garcia, Alexander Romanovsky, Jaelson Castro, and Paulo Alencar, editors. Software Engineering for Multi-Agent Systems II: Research Issues and Practical Applications, volume 2940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21182-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S453 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2940.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21182-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2940;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96018.

Luck:2009:AOS

[LGS09] Michael Luck and Jorge J. Gomez-Sanz, editors. Agent-Oriented Software Engineering IX: 9th International Workshop, AOSE 2008 Estoril, Portugal, May 12–13, 2008 Revised Selected Papers, volume 5386 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01337-6 (print), 3-642-01338-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01338-6.

Lavrac:2003:KDD

[LGTB03] Nada Lavrač, Dragan Gamberger, Ljupčo Todorovski, and Hendrik Blockee, editors. Knowledge Discovery in Databases: PKDD 2003: 7th European Conference on Principles and Practice of Knowledge Discovery in Databases, Cavtat-Dubrovnik, Croatia, September 22–26, 2003: Proceedings, volume 2838 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20085-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 E87 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2838.
htm;   http://www.springerlink.
com/content/978-3-540-
20085-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2838;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13634.

Liu:2006:FMS

[LH06] Zhiming Liu and Jifeng He, editors. Formal Methods and Software Engineering: 8th International Conference on Formal Engineering Methods, ICFEM 2006, Macao, China, November 1–3, 2006. Proceedings, volume 4260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47460-9 (print), 3-540-47462-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        47462-3.

Lopez:2008:CII

[LH08] Javier Lopez and Bernhard M. Hämmerli, editors. Critical Information Infrastructures Security: Second International Workshop, CRITIS 2007, Málaga, Spain, October 3–5, 2007. Revised Papers, volume 5141 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89095-5 (print), 3-540-89173-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89173-4.

Lee:2010:ISC

[LH10] Donghoon Lee and Seokhie Hong, editors. Information, Security and Cryptology — ICISC 2009: 12th International Conference, Seoul, Korea, December 2–4, 2009, Revised Selected Papers, volume 5984 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14422-5 (print), 3-642-14423-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14423-3.

Laaksonen:2011:ASO

[LH11] Jorma Laaksonen and Timo Honkela, editors. Advances in Self-Organizing Maps: 8th International Workshop, WSOM 2011, Espoo, Finland, June 13– 15, 2011. Proceedings, volume 6731 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21565-3 (print), 3-642-21566-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21566-7.

Li:2008:WAS

[LHDD08] Yingshu Li, Dung T. Huynh, Sajal K. Das, and Ding-Zhu Du, editors. Wireless Algorithms, Systems, and Applications: Third International Conference, WASA 2008, Dallas, TX, USA, October 26– 28, 2008. Proceedings, volume 5258 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88581-1 (print), 3-540-88582-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88582-5.

Leitner:2010:HWL

[LHH10] Gerhard Leitner, Martin Hitz, and Andreas Holzinger, editors. HCI in Work and Learning, Life and Leisure: 6th Symposium of the Workgroup Human-Computer Interaction and Usability Engineering, USAB 2010, Klagenfurt, Austria, November 4–5, 2010. Proceedings, volume 6389 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16606-7 (print), 3-642-16607-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16607-5.

Lyons:2011:PCI

[LHH11] Kent Lyons, Jeffrey Hightower, and Elaine M. Huang, editors. Pervasive Computing: 9th International Conference, Pervasive 2011, San Francisco, USA, June 12–15, 2011. Proceedings, volume 6696 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21725-7 (print), 3-642-21726-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21726-5.

Lovrek:2008:KBIa

[LHJ08a] Ignac Lovrek, Robert J.

Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 12th International Conference, KES 2008, Zagreb, Croatia, September 3–5, 2008, Proceedings, Part I, volume 5177 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85562-9 (print), 3-540-85563-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85563-7.

Lovrek:2008:KBIb

[LHJ08b] Ignac Lovrek, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 12th International Conference, KES 2008, Zagreb, Croatia, September 3–5, 2008, Proceedings, Part II, volume 5178 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85564-5 (print), 3-540-85565-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85565-1.

Lovrek:2008:KBIc

[LHJ08c] Ignac Lovrek, Robert J.

Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 12th International Conference, KES 2008, Zagreb, Croatia, September 3–5, 2008, Proceedings, Part III, volume 5179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85566-1 (print), 3-540-85567-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85567-5.

Lopresti:2002:DAS

[LHK02] Daniel Philip Lopresti, Jianying Hu, and Ramanujan Kashi, editors. Document Analysis Systems V: 5th International Workshop, DAS 2002, Princeton, NY, USA, August 19–21, 2002. Proceedings, volume 2423 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44068-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1650 .D37 2002; TK7882.P3 I55 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2423.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44068-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2423.

Lee:2003:MCC

[LhK03] Jaiyong Lee and Chul hee Kang, editors. Mobile communications: 7th CDMA International Conference, CIC 2002, Seoul, Korea, October 29–November 1, 2002: Revised Papers, volume 2524 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00732-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK6570.M6 C4 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2524.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        00732-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2524.

Lee:2012:CHI

[LHKŚ12] Geuk Lee, Daniel Howard, Jeong Jin Kang, and Dominik Ślezak, editors. Convergence and Hybrid Information Technology: 6th International Conference, ICHIT 2012, Daejeon, Korea, August 23–25, 2012. Proceedings, volume 7425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32644-7 (print), 3-642-32645-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32645-5.

Lee:2011:CHI

[LHŚ11] Geuk Lee, Daniel Howard, and Dominik Ślezak, editors. Convergence and Hybrid Information Technology: 5th International Conference, ICHIT 2011, Daejeon, Korea, September 22– 24, 2011. Proceedings, volume 6935 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24081-X (print), 3-642-24082-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24082-9.

Li:1996:RDC

[Li96] S. (Stan) Z. Li, editor. Recent developments in computer vision: Second Asian Conference on Computer Vision, ACCV ’95, Singapore, December 5–8, 1995: invited session papers, volume 1035 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60793-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .A85 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1035.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60793-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1035.

Li:1998:LCP

[Li98] Zhiyuan Li, editor. Languages and compilers for parallel computing: 10th International Workshop, LCPC ’97, Minneapolis, Minnesota, USA August 7–9, 1997: proceedings, volume 1366 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64472-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1366. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1366.htm;
 http://www.springerlink.
com/content/978-3-540-
64472-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1366.

Li:2011:AAA

[Li11a] Jiuyong Li, editor. AI 2010: Advances in Artificial Intelligence: 23rd Australasian Joint Conference, Adelaide, Australia, December 7–10, 2010. Proceedings, volume 6464 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17431-0 (print), 3-642-17432-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17432-2.

Li:2011:DAS

[Li11b] Yingjiu Li, editor. Data and Applications Security and Privacy XXV: 25th Annual IFIP WG 11.3 Conference, DBSec 2011, Richmond, VA, USA, July 11–13, 2011. Proceedings, volume 6818 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22347-8 (print), 3-642-22348-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22348-8.

Liblit:2007:CBI

[Lib07] Ben Liblit, editor. Cooperative Bug Isolation: Winning Thesis of the 2005 ACM Doctoral Dissertation Competition, volume 4440 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71877-X (print), 3-540-71878-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 96 (est.) pp. LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
71878-9.

Lungarella:2007:YAI

[LIBP07] Max Lungarella, Fumiya Iida, Josh Bongard, and Rolf Pfeifer, editors. 50 Years of Artificial Intelligence: Essays Dedicated to the 50th Anniversary of Artificial Intelligence, volume 4850 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77295-2 (print), 3-540-77296-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77296-5.

Lysaght:1999:FPL

[LIH99] Patrick Lysaght, James Irvine, and Reiner Hartenstein, editors. Field programmable logic and applications: 9th International Workshop, FPL ’99, Glasgow, UK, August 30–September 1, 1999: Proceedings, volume 1673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66457-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1673; TK7895.G36 .I48 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1673.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66457-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1673.

Leong:1997:ACI

[LIJ97] Hon Wai Leong, H. Imai, and Sanjay Jain, editors. Algorithms and computation: 8th International Symposium, ISAAC ’97, Singapore, December 17–19, 1997: proceedings, volume 1350 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63890-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1350.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1350.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63890-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1350.

Lind:2001:ISE

[Lin01]  Jurgen Lind.  Itera-

tive Software Engineering for Multiagent Systems: The MASSIVE Method, volume 1994 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42166-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 286 pp. LCCN QA76.758 .L52 2001; QA267.A1 L43 no.1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1994.htm;
 http://www.springerlink.
com/content/978-3-540-
42166-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1994.

Lindell:2003:CSM

[Lin03] Yehuda Lindell. Composition of Secure Multi-Party Protocols: a comprehensive study, volume 2815 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20105-X. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 192 pp. LCCN QA76.9.A25 L57 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2815.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20105-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2815;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13246.

Lin:2007:CCA

[Lin07] Guohui Lin, editor. Computing and Combinatorics: 13th Annual International Conference, COCOON 2007, Banff, Canada, July 16– 19, 2007. Proceedings, volume 4598 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73544-5 (print), 3-540-73545-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73545-8.

Lin:2012:COA

[Lin12] Guohui Lin, editor. Combinatorial Optimization and Applications: 6th International Conference, COCOA 2012, Banff, AB, Canada, August 5– 9, 2012. Proceedings, volume 7402 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31769-3 (print), 3-642-31770-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31770-5.

Lindell:2014:TCT

[Lin14] Yehuda Lindell, editor. Theory of cryptography: 11th Theory of Cryptography Conference, TCC 2014, San Diego, CA, USA, February 24–26, 2014 proceedings, volume 8349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2014. ISBN 3-642-54241-7 (paperback), 3-642-54242-5 (ebk.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 T473 2014. URL http:

//www.springerlink.
com/content/978-3-642-
54242-8.

Liotta:2004:GDI

[Lio04] Giuseppe Liotta, editor. Graph Drawing: 11th International Symposium, GD 2003, Perugia, Italy, September 21–24, 2003: Revised Papers, volume 2912 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20831-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2912.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20831-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2912;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b94919.

Liskov:1981:CRM

[Lis81] B. Liskov. CLU reference manual, volume 114 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10836-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 190 pp. LCCN QA76.73.C23 .C5. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0114.htm;
                                         http://www.springerlink.

com/content/978-0-387-
10836-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=114.

Lisper:1989:SSS

[Lis89] Bjorn Lisper. Synthesizing synchronous systems by static scheduling in space-time, volume 362 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51156-3. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 262 pp. LCCN QA76.5 .L571 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0362.htm;
http://www.springerlink.
com/content/978-0-387-
51156-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=362.

Liu:2001:ESB

[Liu01] Yong Liu, editor. Evolvable systems: from biology to hardware: 4th International Conference, ICES 2001, Tokyo, Japan, October 3–5, 2001: Proceedings, volume 2210 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42671-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .I57 2001; QA267.A1 L43 no.2210. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2210.htm;
 http://www.springerlink.
com/content/978-3-540-
42671-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2210.

Liu:2011:SQA

[Liu11] Weiru Liu, editor. Symbolic and Quantitative Approaches to Reasoning with Uncertainty: 11th European Conference, ECSQARU 2011, Belfast, UK, June 29– July 1, 2011. Proceedings, volume 6717 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22151-3 (print), 3-642-22152-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22152-1.

Lings:2000:ADB

[LJ00] B. J. Lings and Keith Jeffery, editors. Advances in databases: 17th British National Conference on Databases, BN-COD 17, Exeter, UK, July 3–5, 2000: Proceedings, volume 1832 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67743-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1832.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1832.
htm;   http://www.springerlink.
com/content/978-3-540-
67743-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1832.

Li:2007:NPC

[LJJG07] Keqiu Li, Chris Jesshope, Hai Jin, and Jean-Luc Gaudiot, editors. Network and Parallel Computing: IFIP International Conference, NPC 2007, Dalian, China, September 18–21, 2007. Proceedings, volume 4672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74783-4 (print), 3-540-74784-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74784-0.

Lalmas:2010:RAT

[LJR^{+}10] Mounia Lalmas, Joemon Jose, Andreas Rauber, Fabrizio Sebastiani, and Ingo Frommholz, editors. Research and Advanced Technology for Digital Libraries: 14th European Conference, ECDL 2010, Glasgow, UK, September 6–10, 2010. Proceedings, volume 6273 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15463-8 (print), 3-642-15464-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15464-5.

Labarta:2008:HPC

[LJS08] Jesús Labarta, Kazuki Joe, and Toshinori Sato, editors. High-Performance Computing: 6th International Symposium, ISHPC 2005, Nara, Japan, September 7– 9, 2005, First International Workshop on Advanced Low Power Systems, ALPS 2006, Revised Selected Papers, volume 4759 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77703-2 (print), 3-540-77704-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77704-5.

Li:2010:LSMc

[LJS^{+}10] Kang Li, Li Jia, Xin Sun, Minrui Fei, and George W. Irwin, editors. Life System Modeling and Intelligent Computing: International Conference on Life System Modeling and Simulation, LSMS 2010, and International Conference on Intelligent Computing for Sustainable Energy and Environment, IC-SEE 2010, Wuxi, China, September 17–20, 2010. Proceedings, Part III, volume 6330 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15614-2 (print), 3-642-15615-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15615-1.

Laurence:1999:CHD

[LK99] Pierre Laurence and Thomas Kropf, editors. Correct Hardware Design and Verification Methods: 10th IFIP WG10.5 Advanced Research Working Conference, CHARME’99, Bad Herrenalb, Germany, September 27–29, 1999: Proceedings, volume 1703 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66559-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.75 .C453 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1703.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66559-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1703.

Lynch:2000:HSC

[LK00] Nancy A. Lynch and Bruce H. Krogh, editors. Hybrid systems: computation and control: Third International Workshop, HSCC 2000, Pittsburgh, PA, USA, March 223-25, 2000: Proceedings, volume 1790 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67259-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1790. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1790.htm;
 http://www.springerlink.
com/content/978-3-540-
67259-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1790.

Llados:2004:GRR

[LK04] Josep Lladós and Young-Bin Kwon, editors. Graphics Recognition: Recent Advances and Perspectives: 5th International Workshop, GREC 2003, Barcelona, Spain, July 30–31, 2003: Revised Selected Papers, volume 3088 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22478-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .I5538 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3088.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22478-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3088;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99011.

Loffler:2007:MRF

[LK07] Jobst Löffler and Markus Klann, editors. Mobile Response: First International Workshop on Mobile Information Technology for Emergency Response, Mobile Response 2007, Sankt Augustin, Germany, February 22–23, 2007, Revised Selected Papers, volume 4458 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75667-1 (print), 3-540-75668-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75668-2.

Lalanne:2009:HMI

[LK09a] Denis Lalanne and Jürg Kohlas, editors. Human Machine Interaction: Research Results of the MMI Program, volume 5440 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00436-9 (print), 3-642-00437-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00437-7.

Loffler:2009:MRS

[LK09b] Jobst Löffler and Markus Klann, editors. Mobile Response: Second International Workshop on Mobile Information Technology for Emergency Response, MobileResponse 2008. Bonn, Germany, May 29–30, 2008, Revised Selected Papers, volume 5424 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00439-3 (print), 3-642-00440-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00440-7.

Lambrix:2010:DIL

[LK10] Patrick Lambrix and Graham Kemp, editors. Data Integration in the Life Sciences: 7th International Conference, DILS 2010, Gothenburg, Sweden, August 25–27, 2010. Proceedings, volume 6254 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15119-1 (print), 3-642-15120-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15120-0.

Lingas:1993:ALP

[LKC93] Andrzej Lingas, R. Karlsson, and Svante Carlsson, editors. Automata, languages and programming: 20th international colloquium, ICALP 93, Lund, Sweden, July 5–9, 1993: proceedings, volume 700 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56939-1 (Berlin), 0-387-56939-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA267.A1 L43 no.700. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0700.htm;
http://www.springerlink.
com/content/978-3-540-
56939-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=700.

Lastovetsky:2008:RAP

[LKD08] Alexey Lastovetsky, Tahar Kechadi, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 15th European PVM/MPI Users’ Group Meeting, Dublin, Ireland, September 7–10, 2008. Proceedings, volume 5205 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87474-7 (print), 3-540-87475-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87475-1.

Lee:2007:ESS

[LKK^{+}07] Yann-Hang Lee, Heung-Nam Kim, Jong Kim, Yongwan Park, Laurence T. Yang, and Sung Won Kim, editors. Embedded Software and Systems: Third International Conference, ICESS 2007, Daegu, Korea, May 14–16, 2007. Proceedings, volume 4523 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72684-5 (print), 3-540-72685-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72685-2.

Lukowicz:2010:SSC

[LKK10] Paul Lukowicz, Kai Kunze, and Gerd Kortuem, editors. Smart Sensing and Context: 5th European Conference, EuroSSC 2010, Passau, Germany, November 14– 16, 2010. Proceedings, volume 6446 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16981-3 (print), 3-642-16982-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16982-3.

Lippmann:2008:RAI

[LKT08] Richard Lippmann, Engin Kirda, and Ari Trachtenberg, editors. Recent Advances in Intrusion Detection: 11th International Symposium, RAID 2008, Cambridge, MA, USA, September 15–17, 2008. Proceedings, volume 5230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87402-X (print), 3-540-87403-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87403-4.

Li:2008:SEL

[LKZ^{+}08] Xiaodong Li, Michael Kirley, Mengjie Zhang, David Green, Vic Ciesielski, Hussein Abbass, Zbigniew Michalewicz, Tim Hendtlass, Kalyanmoy Deb, Kay Chen Tan, Jürgen Branke, and Yuhui Shi, editors. Simulated Evolution and Learning: 7th International Conference, SEAL 2008, Melbourne, Australia, December 7– 10, 2008. Proceedings, volume 5361 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89693-7 (print), 3-540-89694-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89694-4.

Lupanov:2005:SAF

[LKZCS05] Oleg B. Lupanov, Oktay M. Kasim-Zade, Alexander V. Chaskin, and Kathleen Steinhöfel, editors. Stochastic algorithms: foundations and applications: third international symposium, SAGA 2005, Moscow, Russia, October 20– 22, 2005: proceedings, volume 3777 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29498-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .S24 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29498-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3777.

Lewerentz:1995:FDR

[LL95] Claus Lewerentz and Thomas Lindner. Formal development of reactive systems: case study production cell, volume 891 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-58867-1 (Berlin), 0-387-58867-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 394 pp. LCCN QA76.76.D47 F65 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0891.
htm;   http://www.springerlink.
com/content/978-3-540-
58867-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=891.

Lee:2003:ISC

[LL03] Pil Joong Lee and Chae Hoon Lim, editors. Information security and cryptology — ICISC 2002: 5th International Conference, Seoul, Korea, November 28–29, 2002: Revised Papers, volume 2587 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00716-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I32 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2587.
htm;   http://www.springerlink.
com/content/978-3-540-
00716-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2587.

Also available via the World Wide Web.

Lim:2004:ISC

[LL04] Jong In Lim and Dong Hoon Lee, editors. Information Security and Cryptology—ICISC 2003: 6th International Conference, Seoul, Korea, November 27–28, 2003: Revised Papers, volume 2971 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21376-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2971.htm;
 http://www.springerlink.
com/content/978-3-540-
21376-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2971;
 http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b96249.

Liu:2006:GRT

[LL06] Wenyin Liu and Josep Lladós, editors. Graphics Recognition. Ten Years Review and Future Perspectives: 6th International Workshop, GREC 2005, Hong Kong, China, August 25–26, 2005, Revised Selected Papers, volume 3926 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34711-9 (print), 3-540-34712-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34712-5.

Lee:2007:ABI

[LL07] Seong-Whan Lee and Stan Z. Li, editors. Advances in Biometrics: International Conference, ICB 2007, Seoul, Korea, August 27–29, 2007. Proceedings, volume 4642 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74548-3 (print), 3-540-74549-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74549-5.

Leow:2005:IVR

[LLC^{+}05] Wee-Kheng Leow, Michael S. Lew, Tat-Seng Chua, Wei-Ying Ma, Lekha Chaisorn, and Erwin M. Bakker, editors. Image and video retrieval: 4th international conference, CIVR 2005, Singapore, July 20–22, 2005: proceedings, volume 3568 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27858-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .C49 2005eb; TA1637 .C49 2005; TA1637; Internet. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        27858-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3568.

Leung:2009:NIPa

[LLC09a] Chi Sing Leung, Minho Lee, and Jonathan H. Chan, editors. Neural Information Processing: 16th International Conference, ICONIP 2009, Bangkok, Thailand, December 1–5, 2009, Proceedings, Part I, volume 5863 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10676-5 (print), 3-642-10677-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10677-4.

Leung:2009:NIPb

[LLC09b] Chi Sing Leung, Minho Lee, and Jonathan H. Chan, editors. Neural Information Processing: 16th International Conference, ICONIP 2009, Bangkok, Thailand, December 1–5, 2009, Proceedings, Part II, volume 5864 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10682-X (print), 3-642-10684-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10684-2.

Lau:2005:AWB

[LLCL05] Rynson W. H. Lau, Qing Li, Ronnie Cheung, and Wenyin Liu, editors. Advances in Web-based learning: ICWL 2005: 4th International Conference, Hong Kong, China, July 31–August 3, 2005: Proceedings, volume 3583 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27895-8 (paperback), 3-540-27895-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1044.87 .I39 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        27895-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3583.

Li:2007:LSM

[LLIH07] Kang Li, Xin Li, George William Irwin, and Gusen He, editors. Life System Modeling and Simulation: International Conference, LSMS 2007, Shanghai, China, September 14– 17, 2007. Proceedings, volume 4689 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74770-2 (print), 3-540-74771-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74771-0.

Liu:2006:AWB

[LLL06] Wenyin Liu, Qing Li, and Rynson W. H. Lau, editors. Advances in Web Based Learning — ICWL 2006: 5th International Conference, Penang, Malaysia, July 19–21, 2006. Revised Papers, volume 4181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49027-2 (print), 3-540-68509-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68509-8.

Leung:2008:AWB

[LLLL08] Howard Leung, Frederick Li, Rynson Lau, and Qing Li, editors. Advances in Web Based Learning — ICWL 2007: 6th International Conference Edinburgh, UK, August 15–17, 2007 Revised Papers, volume 4823 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78138-2 (print), 3-540-78139-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78139-4.

Li:2008:IRT

[LLM^{+}08] Hang Li, Ting Liu, Wei-Ying Ma, Tetsuya Sakai, Kam-Fai Wong, and Guodong Zhou, editors. Information Retrieval Technology: 4th Asia Information Retrieval Symposium, AIRS 2008, Harbin, China, January 15–18, 2008 Revised Selected Papers, volume 4993 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68633-9 (print), 3-540-68636-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68636-1.

Lipshteyn:2009:GTC

[LLM09] Marina Lipshteyn, Vadim E. Levit, and Ross M. Mc-Connell, editors. Graph Theory, Computational Intelligence and Thought: Essays Dedicated to Martin Charles Golumbic on the Occasion of His 60th Birthday, volume 5420 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02028-3 (print), 3-642-02029-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02029-2.

Lopes:2009:PAI

[LLMR09] Luís Seabra Lopes, Nuno Lau, Pedro Mariano, and Luís M. Rocha, editors. Progress in Artificial Intelligence: 14th Portuguese Conference on Artificial Intelligence, EPIA 2009, Aveiro, Portugal, October 12–15, 2009. Proceedings, volume 5816 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04685-1 (print), 3-642-04686-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        04686-5.

Langendoerfer:2004:WWI

[LLMT04]  Peter  Langendoerfer,

Mingyan Liu, Ibrahim Matta, and Vassilis Tsaoussidis, editors. Wired/ Wireless Internet Communications: Second International Conference, WWIC 2004: Frankfurt (Oder), Germany, February 4–6, 2004: Proceedings, volume 2957 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20954-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.4885 .W85 2004.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2957.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20954-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2957;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95454.

Lloyd:2000:CLC

[Llo00] J. W. Lloyd, editor. Computational logic— CL 2000: First International Conference, London, UK, July 24– 28, 2000: Proceedings, volume 1861 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67797-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.L63 C65 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1861.htm;
 http://www.springerlink.
com/content/978-3-540-
67797-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1861.

Liu:2008:GRR

[LLO08] Wenyin Liu, Josep Lladós, and Jean-Marc Ogier, editors. Graphics Recognition. Recent Advances and New Opportunities: 7th International Workshop, GREC 2007, Curitiba, Brazil, September 20–21, 2007. Selected Papers, volume 5046 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88184-0 (print), 3-540-88188-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88188-9.

Lecomte:1999:LAC

[LLP99] Alain Lecomte, François Lamarche, and Guy Perrier, editors. Logical aspects of computational linguistics: Second International Conference, LACL ’97, Nancy, France, September 22– 24, 1997: Selected Papers, volume 1582 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65751-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1582; P98.5.L63 L33 1997.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1582.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65751-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1582.

Levy:2007:PPF

[LLP^{+}07] Pierre P. Lévy, Bénédicte Le Grand, François Poulet, Michel Soto, Laszlo Darago, Laurent Toubiana, and Jean-François Vibert, editors. Pixelization Paradigm: First Visual Information Expert Workshop, VIEW 2006, Paris, France, April 24–25, 2006, Revised Selected Papers, volume 4370 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71026-4 (print), 3-540-71027-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71027-1.

Lee:2009:FTD

[LLPH09] David Lee, Antónia Lopes, and Arnd Poetzsch-Heffter, editors. Formal Techniques for Distributed Systems: Joint 11th IFIP WG 6.1 International Conference FMOODS 2009 and 29th IFIP WG 6.1 International Conference FORTE 2009, Lisboa, Portugal, June 9–12, 2009. Proceedings, volume 5522 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02137-9 (print), 3-642-02138-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02138-1.

Laender:2000:CMI

[LLS00] Alberto H. F. Laender, Stephen W. Liddle, and Veda Catherine Storey, editors. Conceptual Modeling—ER 2000: 19th International Conference on Conceptual Modeling: Salt Lake City, Utah, USA, October 9–12, 2000. Proceedings, volume 1920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41072-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1920. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1920.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41072-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1920.

Li:2004:ABP

[LLT^{+}04] Stan Z. Li, Jianhuang Lai, Tieniu Tan, Guocan Feng, and Yunhong Wang, editors. Advances in Biometric Person Authentication: 5th Chinese Conference on Biometric Recognition, SINO-BIOMETRICS 2004, Guangzhou, China, December 13–14, 2004: Proceedings, volume 3338 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24029-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.B56 C4 2004. URL http:

//www.springerlink.
com/content/978-3-540-
24029-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3338;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104239.

LaMarca:2007:PCI

[LLT07] Anthony LaMarca, Marc Langheinrich, and Khai N. Truong, editors. Pervasive Computing: 5th International Conference, PERVASIVE 2007, Toronto, Canada, May 13–16, 2007. Proceedings, volume 4480 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72036-7 (print), 3-540-72037-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72037-9.

Liu:2012:SOC

[LLTY12] Chengfei Liu, Heiko Ludwig, Farouk Toumani, and Qi Yu, editors. Service-Oriented Computing: 10th International Conference, IC-SOC 2012, Shanghai, China, November 12– 15, 2012. Proceedings, volume 7636 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34320-1 (print), 3-642-34321-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34321-6.

Lang:2006:KSE

[LLW06] Jérôme Lang, Fangzhen Lin, and Ju Wang, editors. Knowledge Science, Engineering and Management: First International Conference, KSEM 2006, Guilin, China, August 5–8, 2006. Proceedings, volume 4092 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37033-1 (print), 3-540-37035-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37035-2.

Lee:2004:DSA

[LLWL04] YoonJoon Lee, Jianzhong Li, Kyu-Young Whang, and Doheon Lee, editors. Database Systems for Advanced Applications: 9th International Conference, DASFAA 2004, Jeju Island, Korea, March 17–19, 2004: Proceedings, volume 2973 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21047-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5533 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2973.
htm;   http://www.springerlink.
com/content/978-3-540-
21047-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2973;
 http://www.springerlink.

                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95600.

Liu:2009:WIS

[LLWL09] Wenyin Liu, Xiangfeng Luo, Fu Lee Wang, and Jingsheng Lei, editors. Web Information Systems and Mining: International Conference, WISM 2009, Shanghai, China, November 7–8, 2009. Proceedings, volume 5854 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05249-5 (print), 3-642-05250-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05250-7.

Lamma:1993:ELP

[LM93] E. Lamma and P. Mello, editors. Extensions of logic programming: third international workshop, ELP ’92, Bologna, Italy, February 26–28, 1992: proceedings, volume 660 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56454-3 (Berlin), 0-387-56454-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .E47 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0660.htm;
 http://www.springerlink.
com/content/978-3-540-
56454-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=660.

Lin:1996:ACG

[LM96] Ming C. Lin and Dinesh N. Manocha, editors. Applied computational geometry: towards geometric engineering: FCRC ’96 Workshop, WACG ’96, Philadelphia, PA, May 27–28, 1996: Selected Papers, volume 1148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61785-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA448.D38 A635 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1148.
htm;   http://www.springerlink.
com/content/978-3-540-
61785-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=1148.

Anthology selected from contributions to the First ACM Workshop on Applied Computational Geometry.

Li:1997:ALT

[LM97] Ming Li and Akira Maruoka, editors. Algorithmic learning theory: 8th international workshop, ALT ’97, Sendai, Japan, October 6–8, 1997: proceedings, volume 1316 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63577-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A48 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1316.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63577-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1316.

Lamersdorf:1998:TDS

[LM98a] Winfried Lamersdorf and Michael Merz, editors. Trends in distributed systems for electronic commerce: International IFIP/GI Working Conference, TREC’98, Hamburg, Germany, June 3–5, 1998: Proceedings, volume 1402 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64564-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1402. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1402.htm;
 http://www.springerlink.
com/content/978-3-540-
64564-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1402.

Lee:1998:PTA

[LM98b] Hing-Yan Lee and Hiroshi Motoda, editors. PRICAI’98: topics in artificial intelligence: 5th Pacific Rim International Conference on Artificial Intelligence, Singapore, November 22– 27, 1998: Proceedings, volume 1531 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65271-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P33 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1531.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65271-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1531.

Lucchesi:1998:LTI

[LM98c] Claudio L. Lucchesi and Arnaldo Moura, editors. Latin’98, theoretical informatics: Third Latin American Symposium, Campinas, Brazil, April 1998: Proceedings, volume 1380 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64275-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5.L348 1998. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1380.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        64275-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1380.

Lamma:2000:AAA

[LM00] Evelina Lamma and Paola Mello, editors. AI*IA 99: advances in artificial intelligence: 6th Congress of the Italian Association for Artificial Intelligence, Bologna, Italy, September 14– 17, 1999: Selected Papers, volume 1792 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67350-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1792.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1792.
htm;   http://www.springerlink.
com/content/978-3-540-
67350-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1792.

Lioy:2003:CMS

[LM03] Antonio Lioy and Daniele Mazzocchi, editors. Communications and Multimedia Security: Advanced Techniques for Network and Data Protection: 7th IFIP-TC-11 International Conference, CMS 2003, Torino, Italy, October 2–3, 2003:

Proceedings, volume 2828 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20185-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.59 .C577 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2828.
htm;   http://www.springerlink.
com/content/978-3-540-
20185-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2828;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13863.

Lodaya:2004:FFS

[LM04] Kamal Lodaya and Meena Mahajan, editors. FSTTCS 2004: Foundations of Software Technology and Theoretical Computer Science: 24th International Conference, Chennai, India, December 16–18, 2004. Proceedings, volume 3328 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24058-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
24058-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3328;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104325.

Lamontagne:2006:AAI

[LM06a] Luc Lamontagne and Marion Marchand, editors. Advances in Artificial Intelligence: 19th Conference of the Canadian Society for Computational Studies of Intelligence, Canadian AI 2006, Quebec City, Quebec, Canada, June 7– 9, 2006, Proceedings, volume 4013 of Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34628-7 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34628-9.

Leitold:2006:CMS

[LM06b] Herbert Leitold and Evangelos P. Markatos, editors. Communications and Multimedia Security: 10th IFIP TC-6 TC-11 International Conference, CMS 2006, Heraklion, Crete, Greece, October 19–21, 2006. Proceedings, volume 4237 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47820-5 (print), 3-540-47823-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47823-2.

Leue:2008:FMI

[LM08a] Stefan Leue and Pedro Merino, editors. Formal Methods for Industrial Critical Systems: 12th International Workshop, FMICS 2007, Berlin, Germany, July 1–2, 2007, Revised Selected Papers, volume 4916 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79706-8 (print), 3-540-79707-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79707-4.

Ludascher:2008:SSD

[LM08b] Bertram Ludäscher and Nikos Mamoulis, editors. Scientific and Statistical Database Management: 20th International Conference, SS-DBM 2008, Hong Kong, China, July 9–11, 2008 Proceedings, volume 5069 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69476-5 (print), 3-540-69497-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69497-7.

Leucker:2009:TAC

[LM09] Martin Leucker and Carroll Morgan, editors. Theoretical Aspects of Computing — IC-TAC 2009: 6th International Colloquium, Kuala Lumpur, Malaysia, August 16–20, 2009. Proceedings, volume 5684 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03465-9 (print), 3-642-03466-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03466-4.

Liu:2008:FMS

[LMA08] Shaoying Liu, Tom Maibaum, and Keijiro Araki, editors. Formal Methods and Software Engineering: 10th International Conference on Formal Engineering Methods, ICFEM 2008, Kitakyushu-City, Japan, October 27–31, 2008. Proceedings, volume 5256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88193-X (print), 3-540-88194-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88194-0.

Li:2009:CPO

[LMA09] Wenjie Li and Diego Mollá-Aliod, editors. Computer Processing of Oriental Languages. Language Technology for the Knowledge-based Economy: 22nd International Conference, IC-CPOL 2009, Hong Kong, March 26–27, 2009. Proceedings, volume 5459 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00830-5 (print), 3-642-00831-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00831-3.

Luque:2008:EPP

[LMB08] Emilio Luque, Tomàs Margalef, and Domingo Benítez, editors. Euro-Par 2008 — Parallel Processing: 14th International Euro-Par Conference, Las Palmas de Gran Canaria, Spain, August 26–29, 2008. Proceedings, volume 5168 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85450-9 (print), 3-540-85451-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85451-7.

Liu:2012:ICA

[LMC12] Baoxiang Liu, Maode Ma, and Jincai Chang, editors. Information Computing and Applications: Third International Conference, ICICA 2012, Chengde, China, September 14– 16, 2012. Proceedings, volume 7473 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34061-X (print), 3-642-34062-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34062-8.

Linte:2012:AEC

[LMCH12] Cristian A. Linte, John T. Moore, Elvis C. S. Chen, and David R. Holmes III, editors. Augmented Environments for Computer-Assisted Interventions: 6th International Workshop, AE-CAI 2011, Held in Conjunction with MIC-CAI 2011, Toronto, ON, Canada, September 22, 2011, Revised Selected Papers, volume 7264 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32629-3 (print), 3-642-32630-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32630-1.

Lu:2010:PDS

[LMM10] Chenyang Lu, Toshimitsu Masuzawa, and Mohamed Mosbah, editors. Principles of Distributed Systems: 14th International Conference, OPODIS 2010, Tozeur, Tunisia, December 14–17, 2010. Proceedings, volume 6490 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17652-6 (print), 3-642-17653-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17653-1.

Lee:2010:SDS

[LMMN10] Gary Geunbae Lee, Joseph Mariani, Wolfgang Minker, and Satoshi Nakamura, editors. Spoken Dialogue Systems for Ambient Environments: Second International Workshop on Spoken Dialogue Systems Technology, IWSDS 2010, Gotemba, Shizuoka, Japan, October 1–2, 2010. Proceedings, volume 6392 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16201-0 (print), 3-642-16202-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16202-2.

Lindemann:2004:RAB

[LMP04]  Gabriela  Lindemann,

Daniel Moldt, and Mario Paolucci, editors. Regulated Agent-Based Social Systems: First International Workshop, RASTA 2002, Bologna, Italy, July 16, 2002: Revised Selected and Invited Papers, volume 2934 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20923-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HM701 .R37 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2934.htm;
 http://www.springerlink.
com/content/978-3-540-
20923-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2934;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95469.

Lalmas:2006:AIR

[LMR^{+}06] Mounia Lalmas, Andy MacFarlane, Stefan Rüger, Anastasios Tombros, Theodora Tsikrika, and Alexei Yavlinsky, editors. Advances in Information Retrieval: 28th European Conference on IR Research, ECIR 2006, London, UK, April 10– 12, 2006. Proceedings, volume 3936 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33347-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33347-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3936.

Launchbury:1996:AFP

[LMS96] John Launchbury, Erik Meijer, and T. (Tim) Sheard, editors. Advanced functional programming: second international school, Olympia, WA, USA, August 26– 30, 1996: tutorial text, volume 1129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61628-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .A38 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1129.htm;
 http://www.springerlink.
com/content/978-3-540-
61628-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1129.

Lectures held at Evergreen College.

Lehner:2007:EPP

[LMSS07] Wolfgang Lehner, Norbert Meyer, Achim Streit, and Craig Stewart, editors. Euro-Par 2006: Parallel Processing: Workshops: Core-GRID 2006, UNICORE Summit 2006, Petascale Computational Biology and Bioinformatics, Dresden, Germany, August 29–September 1, 2006, Revised Selected Papers, volume 4375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72226-2 (print), 3-540-72337-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
72337-0.

Luck:2001:MAS

[LMŠT01] Michael Luck, Vladimír Mařík, Olga Štpánková, and Robert Trappl, editors. Multi-agent systems and applications: 9th ECCAI advanced course, ACAI 2001 and agent link’s 3rd European Agent Systems Summer School, EASSS 2001, Prague, Czech Republic, July 2–13, 2001: Selected Tutorial Papers, volume 2086 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42312-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 E85 2001; QA267.A1 L43 no.2086. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2086.htm;
http://www.springerlink.
com/content/978-3-540-
42312-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2086.

Liddle:2000:CMB

[LMT00] Stephen W. Liddle, H. C.

Mayr, and B. Thalheim, editors. Conceptual modeling for Ebusiness and the Web: ER 2000, Workshops on Conceptual Modeling Approaches for E-Business and The World Wide Web and Conceptual Modeling, Salt Lake City, Utah, USA, October 9–12, 2000: Proceedings, volume 1921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41073-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26 I5673 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1921.htm;
 http://www.springerlink.
com/content/978-3-540-
41073-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1921.

Lindner:2004:CTD

[LMT^{+}04] Wolfgang Lindner, Marco Mesiti, Can Türker, Yannis Tzitzikas, and Athena Vakali, editors. Current trends in database technology: EDBT 2004 workshop: EDBT 2004 workshops PHD, Datax, PIM, P2P and DB, and Clustweb, Heraklion, Crete, Greece, March 14– 18, 2004. revised selected papers, volume 3268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23305-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
23305-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3268;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101218.

Lodi:2010:IAT

[LMT10] Andrea Lodi, Michela Milano, and Paolo Toth, editors. Integration of AI and OR Techniques in Constraint Programming for Combinatorial Optimization Problems: 7th International Conference, CPAIOR 2010, Bologna, Italy, June 14– 18, 2010. Proceedings, volume 6140 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13519-6 (print), 3-642-13520-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13520-0.

Ling:1995:DOO

[LMV95] Tok Wang Ling, Alberto O. Mendelzon, and Laurent Vieille, editors. Deductive and objectoriented databases: fourth international conference, DOOD ’95, Singapore, December 4–7, 1995: proceedings, volume 1013 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60608-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55826 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1013.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60608-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1013.

Litwin:1998:ADI

[LMV98] W. Litwin, Tadeusz Morzy, and Gottfried Vossen, editors. Advances in databases and information systems: second East European symposium, ADBIS ’98, Poznan, Poland, September 7–10, 1998: Proceedings, volume 1475 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64924-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1475.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1475.
htm;   http://www.springerlink.
com/content/978-3-540-
64924-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1475.

Lee:2001:RAI

[LMW01] Wenke Lee, Ludovic Mé, and Andreas Wespi, editors. Recent advances in intrusion detection: 4th International Symposium, RAID 2001, Davis, CA, USA, October 10–12, 2001: Proceedings, volume 2212 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42702-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 R34 2001; QA267.A1 L43 no.2212. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2212.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42702-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2212.

Lirkov:2006:LSS

[LMW06] Ivan Lirkov, Svetozar Margenov, and Jerzy Waśniewski, editors. Largescale scientific computing: 5th international conference, lSSC 2005, Sozopol, Bulgaria, June 6–10, 2005, revised papers, volume 3743 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31994-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        31994-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3743.

Lirkov:2008:LSS

[LMW08] Ivan Lirkov, Svetozar Margenov, and Jerzy Waśniewski, editors. Large-Scale Scientific Computing: 6th International Conference, LSSC 2007, Sozopol, Bulgaria, June 5–9, 2007. Revised Papers, volume 4818 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78825-5 (print), 3-540-78827-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78827-0.

Lirkov:2010:LSS

[LMW10] Ivan Lirkov, Svetozar Margenov, and Jerzy Waśniewski, editors. Large-Scale Scientific Computing: 7th International Conference, LSSC 2009, Sozopol, Bulgaria, June 4–8, 2009. Revised Papers, volume 5910 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12534-4 (print), 3-642-12535-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12535-5.

Lirkov:2012:LSS

[LMW12] Ivan Lirkov, Svetozar Margenov, and Jerzy Waśniewski, editors. Large-Scale Scientific Computing: 8th International Conference, LSSC 2011, Sozopol, Bulgaria, June 6–10, 2011, Revised Selected Papers, volume 7116 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29842-7 (print), 3-642-29843-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29843-1.

Lirkov:2004:LSS

[LMWY04] Ivan Lirkov, Svetozar Margenov, Jerzy Waśniewski, and Plamen Yalamov, editors. Large-Scale Scientific Computing: 4th International Conference, LSSC 2003, Sozopol, Bulgaria, June 4– 8, 2003: Revised Papers, volume 2907 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21090-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q183.9 .I33 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2907.htm;

http://www.springerlink.
com/content/978-3-540-
21090-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2907;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95284.

Lakemeyer:1994:FKR

[LN94] Gerhard Lakemeyer and Bernhard Nebel. Foundations of knowledge representation and reasoning, volume 810 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58107-3 (New York), 0-387-58107-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 355 pp. LCCN Q387 .F68 1994. DM72.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0810.
htm;   http://www.springerlink.
com/content/978-3-540-
58107-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=810.

Larsen:2001:CCT

[LN01a] K. G. Larsen and M. Nielsen, editors. CONCUR 2001. concurrency theory: 12th International Conference, Aalborg, Denmark, August 20–25, 2001: Proceedings, volume 2154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42497-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2154.htm;
 http://www.springerlink.
com/content/978-3-540-
42497-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2154.

Little:2001:EHC

[LN01b] M. Reed Little and L. Nigay, editors. Engineering for Human-Computer Interaction: 8th IFIP International Conference, EHCI 2001, Toronto, Canada, May 11–13, 2001: Revised Papers, volume 2254 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-43044-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I35 2001.

URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t2254.
htm;   http://www.springerlink.
com/content/978-3-540-
43044-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2254.

Lifschitz:2003:LPN

[LN03a] Vladimir Lifschitz and Ilkka Niemelä, editors. Logic Programming and Nonmonotonic Reasoning: 7th International Conference, LPNMR 2004, Fort Lauderdale, FL, USA, January 6– 8, 2004: Proceedings, volume 2923 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20721-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .L74 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2923.htm;
 http://www.springerlink.
com/content/978-3-540-
20721-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2923;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94792.

Lingas:2003:FCT

[LN03b] Andrzej Lingas and Bengt J. Nilsson, editors. Fundamentals of Computation Theory: 14th International Symposium, FCT 2003, Malmö, Sweden, August 2003: Proceedings, volume 2751 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40543-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .F367 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2751.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40543-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2751;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11926.

Larsen:2004:FMA

[LN04a] Kim G. Larsen and Peter Niebert, editors. Formal Modeling and Analysis of Timed Systems: First International Workshop, FORMATS 2003, Marseille, France, September 6–7, 2003: Revised Papers, volume 2791 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21671-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2791.htm;
 http://www.springerlink.
com/content/978-3-540-
21671-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2791;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12025.

Lomuscio:2004:DLC

[LN04b] Alessio Lomuscio and Donald Nute, editors. Deontic Logic in Computer Science: 7th International Workshop on Deontic Logic in Computer Science, DEON 2004, Madeira, Portugal, May 26–28, 2004: Proceedings, volume 3065 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22111-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I645 2004. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t3065.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22111-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3065;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98159.

Lee:2009:STE

[LN09] Sunggu Lee and Priya Narasimhan, editors. Software Technologies for Embedded and Ubiquitous Systems: 7th IFIP WG 10.2 International Workshop, SEUS 2009 Newport Beach, CA, USA, November 16– 18, 2009 Proceedings, volume 5860 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10264-6 (print), 3-642-10265-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10265-3.

Leser:2006:DIL

[LNE06] Ulf Leser, Felix Naumann, and Barbara Eckman, editors. Data Integration in the Life Sciences: Third International Workshop, DILS 2006, Hinxton, UK, July 20–22, 2006. Proceedings, volume 4075 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36593-1 (print), 3-540-36595-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36595-2.

Larsen:2006:MICa

[LNS06a] Rasmus Larsen, Mads Nielsen, and Jon Sporring, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2006: 9th International Conference, Copenhagen, Denmark, October 1–6, 2006. Proceedings, Part I, volume 4190 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44707-5 (print), 3-540-44708-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
44708-5.

Larsen:2006:MICb

[LNS06b] Rasmus Larsen, Mads Nielsen, and Jon Sporring, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2006: 9th International Conference, Copenhagen, Denmark, October 1–6, 2006. Proceedings, Part II, volume 4191 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44727-X (print), 3-540-44728-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        44728-3.

Lio:2011:AIS

[LNS11] Pietro Liò, Giuseppe Nicosia, and Thomas Stibor, editors. Artificial Immune Systems: 10th International Conference, ICARIS 2011, Cambridge, UK, July 18–21, 2011. Proceedings, volume 6825 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22370-2 (print), 3-642-22371-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22371-6.

Lowe:2011:MCC

[LNSS11] Benedikt Löwe, Dag Normann, Ivan Soskov, and Alexandra Soskova, editors. Models of Computation in Context: 7th Conference on Computability in Europe, CiE 2011, Sofia, Bulgaria, June 27 — July 2, 2011. Proceedings, volume 6735 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21874-1 (print), 3-642-21875-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21875-0.

Li:2012:RSK

[LNW^{+}12] Tianrui Li, Hung Son Nguyen, Guoyin Wang, Jerzy Grzymala-Busse, Ryszard Janicki, Aboul Ella Hassanien, and Hong Yu, editors. Rough Sets and Knowledge Technology: 7th International Conference, RSKT 2012, Chengdu, China, August 17–20, 2012. Proceedings, volume 7414 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31899-1 (print), 3-642-31900-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31900-6.

Lusk:1988:ICA

[LO88] Ewing Lusk and Ross A. Overbeek, editors. 9th International Conference on Automated Deduction: Argonne, Illinois, USA, May 23–26, 1988: proceedings, volume 310 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19343-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I571 1988. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0310.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        19343-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=310.

Leroy:1998:TCS

[LO98] Xavier Leroy and Atsushi Ohori, editors. Types in compilation: Second International Workshop, TIC’98: Kyoto, Japan, March 25–27, 1998: Proceedings, volume 1473 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64925-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 T52 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1473.
htm;   http://www.springerlink.
com/content/978-3-540-
64925-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1473.

Laender:2002:SPI

[LO02] Alberto H. F. Laender and Arlindo L. Oliveira, editors. String Processing and Information Retrieval: 9th International Symposium, SPIRE 2002, Lisbon, Portugal, September 11– 13, 2002. Proceedings, volume 2476 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44158-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.T48 I58 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2476.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44158-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2476.

Also available via the

World Wide Web.

Lopez-Ortiz:2010:LTI

[LO10] Alejandro López-Ortiz, editor. LATIN 2010: Theoretical Informatics: 9th Latin American Symposium, Oaxaca, Mexico, April 19–23, 2010. Proceedings, volume 6034 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12199-3 (print), 3-642-12200-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12200-2.

Larrosa:2011:RAC

[LO11] Javier Larrosa and Barry O’Sullivan, editors. Recent Advances in Constraints: 14th Annual ERCIM International Workshop on Constraint Solving and Constraint Logic Programming, CSCLP 2009, Barcelona, Spain, June 15–17, 2009, Revised Selected Papers, volume 6384 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19485-0 (print), 3-642-19486-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19486-3.

Loeckx:1974:ALP

[Loe74] Jacques Loeckx, editor. Automata, languages and programming. Second Colloquium, University of Saarbrucken, July 29–August 2, 1974, volume 14 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 0-387-06841-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.14.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0014.
htm;   http://www.springerlink.
com/content/978-0-387-

                                        06841-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=14.

English, French and German. Sponsored by the Gesellschaft für Informatik and organized in cooperation with the Special Interest Group on Automata and Computability Theory (SIGACT) and with the European Association for Theoretical Computer Science (EATCS).

Lopez-Ortiz:2005:CAA

[LOH05] Alejandro López-Ortiz and Angèle Hamel, editors. Combinatorial and Algorithmic Aspects of Networking: First Workshop on Combinatorial and Algorithmic Aspects of Networking, CAAN 2004, Banff, Alberta, Canada, August 5– 7, 2004, Revised Selected Papers, volume 3405 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27873-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        27873-3;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3405.

Lomet:1993:FDO

[Lom93] David B. Lomet, editor. Foundations of data organization and algorithms: 4th International Conference, FODO ’93, Chicago, Illinois, USA, October 13–15, 1993: proceedings, volume 730 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57301-1, 0-387-57301-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5535 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0730.htm;
 http://www.springerlink.
com/content/978-3-540-
57301-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=730.

Lomas:1997:SPI

[Lom97] Mark Lomas, editor. Security protocols: International Workshop: Cambridge, United Kingdom, April 10–12, 1996, proceedings, volume 1189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62494-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C826 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1189.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62494-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1189.

Long:1990:SEE

[Lon90] F. W. Long, editor. Software engineering environments: proceedings, volume 467 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 3-540-53452-0 (Berlin), 0-387-53452-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I576 1989. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0467.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        53452-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=467.

Lopez:2006:CII

[Lop06] Javier Lopez, editor. Critical Information Infrastructures Security: First International Workshop, CRITIS 2006, Samos, Greece, August 31 — September 1, 2006. Revised Papers, volume 4347 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69083-2 (print), 3-540-69084-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69084-9.

Lorenz:2001:NIFa

[Lor01a] Pascal Lorenz, editor.

Networking—ICN 2001: First International Conference on Networking, Colmar, France, July 9– 13, 2001: proceedings: Part I, volume 2093 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42302-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2093-2094; TK5105.5 .N46695 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2093.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        42302-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2093.

Lorenz:2001:NIFb

[Lor01b] Pascal Lorenz, editor. Networking—ICN 2001: First International Conference on Networking, Colmar, France, July 9– 13, 2001: proceedings: Part II, volume 2094 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42303-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2093-2094; TK5105.5 .N46695 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2094.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42303-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2094.

Leavens:2010:VST

[LOR10] Gary T. Leavens, Peter O’Hearn, and Sriram K. Rajamani, editors. Verified Software: Theories, Tools, Experiments: Third International Conference, VSTTE 2010, Edinburgh, UK, August 16–19, 2010. Proceedings, volume 6217 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15056-X (print), 3-642-15057-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15057-9.

Li:2005:CAG

[LOS05] Hongbo Li, Peter J.

Olver, and Gerald Sommer, editors. Computer Algebra and Geometric Algebra with Applications: 6th International Workshop, IWMM 2004, Shanghai, China, May 19–21, 2004 and International Workshop, GIAE 2004, Xian, China, May 24–28, 2004, Revised Selected Papers, volume 3519 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26296-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 I59 2006.

URL http://www.springerlink.
com/content/978-3-540-
26296-1;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3519.

Leite:2004:DAL

[LOST04] João Leite, Andrea Omicini, Leon Sterling, and Paolo Torroni, editors. Declarative Agent Languages and Technologies: First International Workshop, DALT 2003, Melbourne, Australia, July 15, 2003, Revised Selected and Invited Papers, volume 2990 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22124-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .I5775 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2990.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22124-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2990;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97923.

Leite:2005:DAL

[LOTY05] João Leite, Andrea Omicini, Paolo Torroni, and Pinar Yolum, editors. Declarative Agent Languages and Technologies II: Second International Workshop, DALT 2004, New York, NY, USA, July 19, 2004, Revised Selected Papers, volume 3476 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26172-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .I5775 2004. URL http:

//www.springerlink.
com/content/978-3-540-
26172-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3476.

Loucopoulos:1992:AIS

[Lou92] P. Loucopoulos, editor.

Advanced information systems engineering: 4th International Conference CAiSE ’92, Manchester, UK, May 12–15, 1992: proceedings, volume 593 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55481-5 (Berlin), 0-387-55481-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .N67 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0593.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55481-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=593.

Loucopoulos:1994:ERA

[Lou94] P. Loucopoulos, editor. Entity-relationship approach — ER ’94: business modelling and reengineering: 13th International Conference on the Entity-Relationship Approach, Manchester, United Kingdom, December 13–16, 1994: proceedings, volume 881 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58786-1 (New York), 3-540-58786-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5583 1994. DM104.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0881.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        58786-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=881.

Loveland:1982:CAD

[Lov82] Donald W. Loveland, editor. 6th Conference on Automated Deduction, New York, USA, June 7– 9, 1982, volume 138 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11558-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 .C66 1982. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0138.htm;
 http://www.springerlink.
com/content/978-0-387-
11558-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=138.

Lam:1999:ACA

[LOX99] Kwok Yan Lam, Eiji Okamato, and Chaoping Xing, editors. Advances in cryptology — ASIACRYPT’99: International Conference on the Theory and Application of Cryptology and Information Security, Singapore, November 14–18, 1999: Proceedings, volume 1716 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66666-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I555 1999.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1716.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66666-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1716.

Leake:1997:CBR

[LP97] David B. Leake and Enric Plaza, editors. Casebased reasoning: research and development: Second International Conference, ICCBR-97, Providence, RI, USA, July 25–27, 1997: proceedings, volume 1266 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63233-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95C3762 1997.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1266.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63233-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1266.

LopezdeMantaras:2000:MLE

[LP00] Ramon López de Mántaras and Enric Plaza, editors. Machine learning: ECML 2000: 11th European Conference on Machine Learning, Barcelona, Catalonia, Spain, May 31–June 2, 2000, volume 1810 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67602-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1810.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1810.
htm;   http://www.springerlink.
com/content/978-3-540-
67602-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1810.

Luck:2008:AOS

[LP08] Michael Luck and Lin Padgham, editors. Agent-Oriented Software Engineering VIII: 8th International Workshop, AOSE 2007, Honolulu, HI, USA, May 14, 2007, Revised Selected Papers, volume 4951 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79487-5 (print), 3-540-79488-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79488-2.

Lilius:2010:ATP

[LP10a] Johan Lilius and Wojciech Penczek, editors. Applications and Theory of Petri Nets: 31st International Conference, PETRI NETS 2010, Braga, Portugal, June 21–25, 2010. Proceedings, volume 6128 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13674-5 (print), 3-642-13675-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13675-7.

Link:2010:FIK

[LP10b] Sebastian Link and Henri Prade, editors. Foundations of Information and Knowledge Systems: 6th International Symposium, FoIKS 2010, Sofia, Bulgaria, February 15–19, 2009. Proceedings, volume 5956 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11828-3 (print), 3-642-11829-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11829-6.

Loganantharaj:2000:IPS

[LPA00] Rasiah Loganantharaj, Gunther Palm, and Moonis Ali, editors. Intelligent problem solving: methodologies and approaches: 13th International Conference on Industrial and Engineering Applications of Artificial Intelligence and Expert Systems, IEA/AIE 2000, New Orleans, Louisiana, USA, June 19–22, 2000: Proceedings, volume 1821 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67689-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1821. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1821.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        67689-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1821.

Leung:2011:AWB

[LPC^{+}11] Howard Leung, Elvira Popescu, Yiwei Cao, Rynson W. H. Lau, and Wolfgang Nejdl, editors. Advances in Web-Based Learning — ICWL 2011: 10th International Conference, Hong Kong, China, December 8–10, 2011. Proceedings, volume 7048 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25812-3 (print), 3-642-25813-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25813-8.

Linnhoff-Popien:2000:TDS

[LPH00] Claudia Linnhoff-Popien and Heinz-Gerd Hegering, editors. Trends in distributed systems: towards a universal service market: Third International IFIP/GI Working Conference, USM 2000, Munich, Germany, September 12–14, 2000: Proceedings, volume 1890 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41024-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 U86 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1890.htm;
 http://www.springerlink.
com/content/978-3-540-
41024-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1890.

Lewis:2009:CBS

[LPH09] Grace A. Lewis, Iman Poernomo, and Christine Hofmeister, editors. Component-Based Software Engineering: 12th International Symposium, CBSE 2009 East Stroudsburg, PA, USA, June 24–26, 2009 Proceedings, volume 5582 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02413-0 (print), 3-642-02414-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02414-6.

Lopez:2005:TPS

[LPK05] Javier López, Günther Pernul, and Sokratis K. Katsikas, editors. Trust, Privacy, and Security in Digital Business: Second International Conference, TrustBus 2005, Copenhagen, Denmark, August 22–26, 2005: Proceedings, volume 3592 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28224-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28224-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3592.

Larsen:2006:FQA

[LPOA^{+}06] Henrik Legind Larsen, Gabriella Pasi, Daniel Ortiz-Arroyo, Troels Andreasen, and Henning Christiansen, editors. Flexible Query Answering Systems: 7th International Conference, FQAS 2006, Milan, Italy, June 7–10, 2006 Proceedings, volume 4027 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34638-4 (print), 3-540-34639-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34639-5.

Lodi:2008:IPC

[LPR08] Andrea Lodi, Alessandro Panconesi, and Giovanni Rinaldi, editors. Integer Programming and Combinatorial Optimization: 13th International Conference, IPCO 2008 Bertinoro, Italy, May 26–28, 2008 Proceedings, volume 5035 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68886-2 (print), 3-540-68891-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68891-4.

Li:2012:AHM

[LPR12] Xiang-Yang Li, Symeon Papavassiliou, and Stefan Ruehrup, editors. Ad-hoc, Mobile, and Wireless Networks: 11th International Conference, ADHOC-NOW 2012, Belgrade, Serbia, July 9–11, 2012. Proceedings, volume 7363 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31637-9 (print), 3-642-31638-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31638-8.

Lampson:1983:DSA

[LPSD83] Butler W. Lampson, Manfred Paul, H. J. Siegert, and David William Davies, editors. Distributed systems, architecture and implementation: an advanced course. Springer-Verlag Inc., New York, NY, USA, 1983. CODEN LNCSD9. ISBN 3-540-12116-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D6 1983.

URL http://www.springerlink.
com/content/978-3-540-
12116-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3945.

Originally published as [LSDP81]. Papers presented at the Advanced Course on Distributed Systems — Architecture and Implementation which was organized by the Institute für Informatik, Technische Universität München, and held March 4–13, 1980.

Lacoste:2004:SSE

[LPSW04] Gérard Lacoste, Birgit Pfitzmann, Michael Steiner, and Michael Waidner. SEMPER— Secure Electronic Marketplace for Europe, volume 1854 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-67825-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 350 pp. LCCN HF5548.32 .S46 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1854.
htm;   http://www.springerlink.
com/content/978-3-540-
67825-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1854;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b75215.

Lambrinoudakis:2007:TPS

[LPT07] Costas Lambrinoudakis, Günther Pernul, and A. Min Tjoa, editors. Trust, Privacy and Security in Digital Business:

4th International Conference, TrustBus 2007, Regensburg, Germany, September 3–7, 2007. Proceedings, volume 4657 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74408-8 (print), 3-540-74409-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74409-2.

Logozzo:2008:VMC

[LPZ08] Francesco Logozzo, Doron A. Peled, and Lenore D. Zuck, editors. Verification, Model Checking, and Abstract Interpretation: 9th International Conference, VM-CAI 2008, San Francisco, USA, January 7– 9, 2008. Proceedings, volume 4905 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78162-5 (print), 3-540-78163-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78163-9.

Levine:2012:MPM

[LPZ12] Joshua A. Levine, Rasmus R. Paulsen, and Yongjie Zhang, editors. Mesh Processing in Medical Image Analysis 2012: MICCAI 2012 International Workshop, MeshMed 2012, Nice, France, October 1, 2012. Proceedings, volume 7599 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33462-8 (print), 3-642-33463-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33463-4.

Lopez:2004:ICS

[LQO04] Javier Lopez, Sihan Qing, and Eiji Okamoto, editors. Information and communications security: 6th international conference, ICICS 2004, Malaga, Spain, October 27–29, 2004: proceedings, volume 3269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23563-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.9.A25 I554 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23563-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3269.

Litwin:1994:ADF

[LR94] W. Litwin and Tore Risch, editors. Applications of databases: first international conference, ADB-94, Vadstena, Sweden, June 21–23, 1994: proceedings, volume 819 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58183-9 (Berlin), 0-387-58183-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 A67 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0819.htm;
 http://www.springerlink.
com/content/978-3-540-
58183-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=819.

“The First International Conference on Applications of Databases (ADB-94) was held … in Vadstena, Sweden”– Foreword.

Liskiewicz:2005:FCT

[LR05a] Maciej Liśkiewicz and Rüdiger Reischuk, editors. Fundamentals of Computation Theory: 15th International Symposium, FCT 2005, Lübeck, Germany, August 17–20, 2005: Proceedings, volume 3623 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28193-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .F367 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3623;
http://www.springerlink.
com/content/978-3-540-
28193-1.

Ludascher:2005:DIL

[LR05b] Bertram Ludäscher and Louiqa Raschid, editors. Data Integration in the Life Sciences: Second International Workshop, DILS 2005, San Diego, CA, USA, July 20–22, 2005: Proceedings, volume 3615 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27967-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .D54 2005. URL http:

//www.springerlink.
com/content/978-3-540-
27967-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3615.

Liu:2009:ATV

[LR09] Zhiming Liu and Anders P. Ravn, editors. Automated Technology for Verification and Analysis: 7th International Symposium, ATVA 2009, Macao, China, October 14–16, 2009. Proceedings, volume 5799 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04760-2 (print), 3-642-04761-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04761-9.

Lemaitre:2004:AAI

[LRG04] Christian Lemaître, Carlos A. Reyes, and Jesús A. González, editors. Advances in Artificial Intelligence— IBERAMIA 2004: 9th Ibero–American Conference on AI, Puebla, Mexico, November 22– 25, 2004, Proceedings, volume 3315 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23806-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3315.htm;
 http://www.springerlink.
com/content/978-3-540-
23806-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3315;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b102591.

Loftsson:2010:ANL

[LRH10] Hrafn Loftsson, Eiríkur Rögnvaldsson, and Sigrún Helgadóttir, editors. Advances in Natural Language Processing: 7th International Conference on NLP, IceTAL 2010, Reykjavik, Iceland, August 16–18, 2010, volume 6233 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14769-0 (print), 3-642-14770-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14770-8.

Ling:1998:CMI

[LRL98] Tok Wang Ling, Sudha Ram, and Mong Li Lee, editors. Conceptual modeling — ER ’98: 17th International Conference on Conceptual Modeling, Singapore, November 116-19, 998: Proceedings, volume 1507 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65189-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26I567 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1507.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65189-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1507.

Luby:1998:RAT

[LRS98] Michael Luby, Jose Rolim, and Maria Serna, editors. Randomization and Approximation Techniques in Computer Science: Second International Workshop, RAN-DOM’98, Barcelona, Spain, October 8–10, 1998. Proceedings, volume 1518 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65142-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.R2655 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1518.
htm;   http://www.springerlink.
com/content/978-3-540-
65142-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1518.

Laddaga:2003:SAS

[LRS03] Robert Laddaga, Paul Robertson, and Howard E. Shrobe, editors. Selfadaptive software: applications: Second International Workshop, IWSAS 2001, Balatonfüred, Hungary, May 17–19, 2001: Revised Papers, volume 2614 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00731-8 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.S375 I97 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2614.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00731-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2614.

Lepisto:1988:ALP

[LS88] Timo Lepisto and Arto Salomaa, editors. Automata, languages and programming: 15th International Colloquium, Tampere, Finland, July 11–15, 1988 proceedings, volume 317 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19488-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.317. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0317.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        19488-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=317.

Proceedings of ICALP

88.

Litwin:1989:FDO

[LS89] W. Litwin and H.-J Schek, editors. Foundations of data organization and algorithms: 3rd international conference, FODO 1989, Paris, France, June 21–23, 1989: proceedings, volume 367 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51295-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 F671 1989. US$39.40. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0367.
htm;   http://www.springerlink.
com/content/978-0-387-
51295-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=367.

One contribution in French.

Larsen:1992:CAV

[LS92] Kim G. Larsen and Arne Skou, editors. Computer aided verification: 3rd international workshop, CAV ’91, Aalborg, Denmark, July 1–4, 1991: proceedings, volume 575 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55179-4 (Berlin), 0-387-55179-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E93 C38 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0575.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55179-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=575.

Lee:1995:CCT

[LS95] Insup Lee and Scott A. Smolka, editors. CON-CUR ’95: concurrency theory: 6th International Conference, Philadelphia, PA, USA, August 21–24, 1995: proceedings, volume 962 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60218-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0962.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60218-7;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=962.

Lausen:2004:DPL

[LS04a] Georg Lausen and Dan Suciu, editors. Database Programming Languages: 9th International Workshop, DBPL 2003, Potsdam, Germany, September 6–8, 2003: Revised Papers, volume 2921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20896-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I585 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2921.
htm;   http://www.springerlink.
com/content/978-3-540-
20896-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2921;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95340.

Llamosa:2004:RST

[LS04b] Albert Llamosà and Alfred Strohmeier, editors. Reliable Software Technologies—Ada-Europe 2004: 9th ADA-Europe International Conference on Reliable Software Technologies, Palma de Mallorca, Spain, June 14–18, 2004, Proceedings, volume 3063 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22011-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3063.htm;
 http://www.springerlink.
com/content/978-3-540-
22011-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3063;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97913.

Leue:2005:SMT

[LS05] Stefan Leue and Tarja Johanna Systä, editors. Scenarios: Models, Transformations and Tools: International Workshop, Dagstuhl Castle, Germany, September 7–12, 2003, Revised Selected Papers, volume 3466 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26189-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S3 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26189-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3466.

Lightfoot:2006:MPL

[LS06a] David E. Lightfoot and Clemens Szyperski, editors. Modular Programming Languages: 7th Joint Modular Languages Conference, JMLC 2006 Oxford, UK, September 13–15, 2006 Proceedings, volume 4228 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-40927-0 (print), 3-540-40928-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
40928-1.

Lowe:2006:SCI

[LS06b] Welf Löwe and Mario Südholt, editors. Software Composition: 5th International Symposium, SC 2006 Vienna, Austria, March 25–26, 2006 Revised Papers, volume 4089 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37657-7 (print), 3-540-37659-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37659-0.

Lugosi:2006:LTA

[LS06c] Gábor Lugosi and Hans Ulrich Simon, editors. Learning Theory: 19th Annual Conference on Learning Theory, COLT 2006, Pittsburgh, PA, USA, June 22–25, 2006. Proceedings, volume 4005 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35294-5 (print), 3-540-35296-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        35296-9.

Lukose:2009:MAS

[LS09] Dickson Lukose and Zhongzhi Shi, editors. Multi-Agent Systems for Society: 8th Pacific Rim International Workshop on Multi-Agents, PRIMA 2005, Kuala Lumpur, Malaysia, September 26–28, 2005, Revised Selected Papers, volume 4078 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03337-7 (print), 3-642-03339-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03339-1.

Laneve:2010:WSF

[LS10a] Cosimo Laneve and Jianwen Su, editors. Web Services and Formal Methods: 6th International Workshop, WS-FM 2009, Bologna, Italy, September 4–5, 2009, Revised Selected Papers, volume 6194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14457-8 (print), 3-642-14458-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14458-5.

Lynch:2010:DCI

[LS10b] Nancy A. Lynch and Alexander A. Shvartsman, editors. Distributed Computing: 24th International Symposium, DISC 2010, Cambridge, MA, USA, September 13–15, 2010. Proceedings, volume 6343 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15762-9 (print), 3-642-15763-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15763-9.

Lassiter:2012:NDL

[LS12a] Daniel Lassiter and Marija Slavkovik, editors. New Directions in Logic, Language and Computation: ESSLLI 2010 and ESSLLI 2011 Student Sessions. Selected Papers, volume 7415 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31466-X (print), 3-642-31467-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31467-4.

Lukasiewicz:2012:FIK

[LS12b] Thomas Lukasiewicz and Attila Sali, editors. Foundations of Information and Knowledge Systems: 7th International Symposium, FoIKS 2012, Kiel, Germany, March 5–9, 2012. Proceedings, volume 7153 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28471-X (print), 3-642-28472-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28472-4.

Li:2004:GCCa

[LSDN04a] Minglu Li, Xian-He Sun, Qianni Deng, and Jun Ni, editors. Grid and Cooperative Computing: Second International Workshop, GCC 2003, Shanghai, China, December 7–10, 2003, Revised Papers, Part I, volume 3032 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21988-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3032.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21988-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3032;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97162.

Li:2004:GCCb

[LSDN04b] Minglu Li, Xian-He Sun, Qianni Deng, and Jun Ni, editors. Grid and Cooperative Computing: Second International Workshop, GCC 2003, Shanghai, China, December 7–10, 2003, Revised Papers, Part II, volume 3033 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21993-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3033.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21993-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3033;

http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97163.

Lampson:1981:DSA

[LSDP81] Butler W. Lampson, H. J. Siegert, Donald Watts Davies, and Manfred Paul, editors. Distributed systems — architecture and implementation: an advanced course, volume 105 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10571-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D56. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0105.
htm;   http://www.springerlink.
com/content/978-0-387-
10571-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=105.

Republished in [LPSD83]. Papers presented at the Advanced Course on Distributed Systems — Architecture and Implementation, which was organized by the Institute für Informatik, Technische Universität München, and held March 4–13, 1980. Contents: Motivations, objectives, and characterization of distributed systems / Gerard LeLann — Distributed system architecture model / Richard W. Watson — Interprocess communication layer / Kenneth J. Thurber — Hardware interconnection technology / Kenneth J. Thurber — Link level / Gerard LeLann — Hierarchy / Donald W. Davies, Richard W. Watson — IPC interface and end-to-end protocols / Richard W. Watson — Distributed control / E. Douglas Jensen — Indentifiers (naming) in distributed systems / Richard W. Watson — Protection / Donald W. Davies — Atomic transactions / Butler W. Lampson — Synchronization / Gerard LeLann — Multiple copy update / Elmar Holler — Applications and protocols / Stephen R. Kimbleton, Pearl Wong, and Butler W. Lampson — Error recovery / Gerard LeLann — Hardware issues / Kenneth J. Thurber — Hardware/software relationships in distributed systems / E. Douglas Jensen — The national software works (NSW) / Elmar Holler — Ethernet, Pup, and Violet / Butler W.

Lampson — Conclusion / Kenneth J. Thurber.

Lew:2002:IVR

[LSE02] Michael S. Lew, Nicu Sebe, and J. P. Eakins, editors. Image and Video Retrieval: International Conference, CIVR 2002, London, UK, July 18–19, 2002. Proceedings, volume 2383 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43899-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .C49 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2383.htm;
http://www.springerlink.
com/content/978-3-540-
43899-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2383.

Lopez:2007:PKI

[LSF07] Javier Lopez, Pierangela Samarati, and Josep L. Ferrer, editors. Public Key Infrastructure: 4th European PKI Workshop: Theory and Practice, EuroPKI 2007, Palma de Mallorca, Spain, June 28–30, 2007. Proceedings, volume 4582 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73407-4 (print), 3-540-73408-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73408-6.

Lee:2006:DEI

[LSgL^{+}06] Juhnyoung Lee, Junho Shim, Sang goo Lee, Christoph Bussler, and Simon Shim, editors. Data Engineering Issues in E-Commerce and Services: Second International Workshop, DEECS 2006, San Francisco, CA, USA, June 26, 2006. Proceedings, volume 4055 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35440-9 (print), 3-540-35441-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35441-3.

Lew:2007:HCI

[LSHB07] Michael Lew, Nicu Sebe, Thomas S. Huang, and Erwin M. Bakker, editors. Human–Computer Interaction: IEEE International Workshop, HCI 2007 Rio de Janeiro, Brazil, October 20, 2007 Proceedings, volume 4796 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75772-4 (print), 3-540-75773-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75773-3.

Liu:2011:MBI

[LSIT11] Tianming Liu, Dinggang Shen, Luis Ibanez, and Xiaodong Tao, editors. Multimodal Brain Image Analysis: First International Workshop, MBIA 2011, Held in Conjunction with MICCAI 2011, Toronto, Canada, September 18, 2011. Proceedings, volume 7012 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24445-9 (print), 3-642-24446-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        24446-9.

LaPoutre:2006:AME

[LSJ06] Han La Poutré, Norman M. Sadeh, and Sverker Janson, editors. Agent-Mediated Electronic Commerce. Designing Trading Agents and Mechanisms: AA-MAS 2005 Workshop, AMEC 2005, Utrecht, Netherlands, July 25, 2005, and IJCAI 2005 Workshop, TADA 2005, Edinburgh, UK, August 1, 2005, Selected and Revised Papers, volume 3937 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46242-2 (print), 3-540-46243-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46243-9.

Lipp:2008:TCC

[LSK08] Peter Lipp, Ahmad-Reza Sadeghi, and Klaus-Michael Koch, editors. Trusted Computing — Challenges and Applications: First International Conference on Trusted Computing and Trust in Information Technologies, Trust 2008 Villach, Austria, March 11–12, 2008 Proceedings, volume 4968 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68978-8 (print), 3-540-68979-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68979-9.

Liu:2004:AWB

[LSL04] Wenyin Liu, Yuanchun Shi, and Qing Li, editors. Advances in Web-based learning-ICWL 2004: third international conference, Beijing, China, August 8–11, 2004, proceedings, volume 3143 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22542-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
22542-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3143.

Lee:2009:IRT

[LSL^{+}09] Gary Geunbae Lee, Dawei Song, Chin-Yew Lin, Akiko Aizawa, Kazuko Kuriyama, Masaharu Yoshioka, and Tetsuya Sakai, editors. Information Retrieval Technology: 5th Asia Information Retrieval Symposium, AIRS 2009, Sapporo, Japan, October 21–23, 2009. Proceedings, volume 5839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04768-8 (print), 3-642-04769-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04769-5.

Lawry:2003:MWL

[LSR03] Jonathan Lawry, James G. Shanahan, and Anca L. Ralescu, editors. Modelling with Words: Learning, Fusion, and Reasoning Within a Formal Linguistic Representation Framework, volume 2873 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20487-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN P98 .M63 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2873.htm;
 http://www.springerlink.
com/content/978-3-540-
20487-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2873;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94063.

Lange:2002:DSI

[LSS02] Steffen Lange, Ken Satoh, and Carl H. Smith, editors. Discovery Science: 5th International Conference, DS 2002: Lübeck, Germany, November 24–26, 2002. Proceedings, volume 2534 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00188-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q174 .I5625 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2534.
htm;   http://www.springerlink.
com/content/978-3-540-
00188-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2534.

Liew:2004:PDC

[LSS^{+}04] Kim-Meow Liew, Hong Shen, Simon See, Wentong Cai, Pingzhi Fan, and Susumu Horiguchi, editors. Parallel and distributed computing, applications and technologies: 5th international conference, PD-CAT 2004, Singapore, December 8–10, 2004: proceedings, volume 3320 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24013-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.+. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24013-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3320;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b103538.

Lakemeyer:2007:RRS

[LSST07] Gerhard Lakemeyer, Elizabeth Sklar, Domenico G. Sorrenti, and Tomoichi Takahashi, editors. RoboCup 2006: Robot Soccer World Cup X, volume 4434 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74023-6 (print), 3-540-74024-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74024-7.

Li:2005:ABPb

[LST^{+}05] Stan Z. Li, Zhenan Sun, Tieniu Tan, Sharath Pankanti, Gérard Chollet, and David Zhang, editors. Advances in biometric person authentication: International Workshop on Biometric Recognition Systems, IWBRS 2005, Beijing, China, October 22–23, 2005: proceedings, volume 3781 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29431-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.B56 I58 2005.

URL http://www.springerlink.
com/content/978-3-540-
29431-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3781.

Lones:2012:IPC

[LST^{+}12] Michael A. Lones, Stephen L.

Smith, Sarah Teichmann, Felix Naef, James A. Walker, and Martin A. Trefzer, editors. Information Processing in Cells and Tissues: 9th International Conference, IPCAT 2012, Cambridge, UK, March 31 — April 2, 2012. Proceedings, volume 7223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28791-3 (print), 3-642-28792-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28792-3.

Liddle:2012:DESa

[LSTZ12a] Stephen W. Liddle, Klaus-Dieter Schewe, A. Min Tjoa, and Xiaofang Zhou, editors. Database and Expert Systems Applications: 23rd International Conference, DEXA 2012, Vienna, Austria, September 3–6, 2012. Proceedings, Part I, volume 7446 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32599-8 (print), 3-642-32600-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32600-4.

Liddle:2012:DESb

[LSTZ12b] Stephen W. Liddle, Klaus-Dieter Schewe, A. Min Tjoa, and Xiaofang Zhou, editors. Database and Expert Systems Applications: 23rd International Conference, DEXA 2012, Vienna, Austria, September 3–6, 2012. Proceedings, Part II, volume 7447 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32596-3 (print), 3-642-32597-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32597-7.

Lu:2007:CSJ

[LSU07] Ruqian Lu, Jörg H.

Siekmann, and Carsten Ullrich, editors. Cognitive Systems: Joint Chinese-German Workshop, Shanghai, China, March 7–11, 2005, Revised Selected Papers, volume 4429 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70933-9 (print), 3-540-70934-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70934-3.

Lammel:2006:GTT

[LSV06] Ralf Lämmel, João Saraiva, and Joost Visser, editors. Generative and Transformational Techniques in Software Engineering: International Summer School, GTTSE 2005, Braga, Portugal, July 4–8, 2005. Revised Papers, volume 4143 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45778-X (print), 3-540-46235-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46235-4.

Ledgard:1981:DHF

[LSW81] Henry F. Ledgard, Andrew Singer, and John Whiteside. Directions in human factors for interactive systems, volume 103 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10574-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 190 pp. LCCN QA76.9.I58 .L43. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0103.
htm;   http://www.springerlink.
com/content/978-0-387-
10574-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=103.

Larsen:1998:ALP

[LSW98] K. G. (Kim G.) Larsen, Sven Skyum, and G. Winskel, editors. Automata, languages, and programming: 25th international colloquium, ICALP’98, Aalborg, Denmark, July 13–17, 1998: Proceedings, volume 1443 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64781-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I55 1998. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1443.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64781-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1443.

The International Colloquium on Automata, Languages, and Programming (ICALP) is the Annual Conference series of the European Association for Theoretical Computer Science (EATCS).

Lanzi:2000:LCS

[LSW00] Pier Luca Lanzi, Wolfgang Stolzmann, and Stewart W. Wilson, editors. Learning classifier systems: from foundations to applications: Proceedings of the Second International Workshop on Learning Classifier Systems, held July 13, 1999 in Orlando, Florida, volume 1813 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67729-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .L423 2000. URL http:

                                        //link.springer-ny.

com/link/service/series/
0558/tocs/t1813.htm;
 http://www.springerlink.
com/content/978-3-540-
67729-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1813.

Lanzi:2001:ALC

[LSW01] Pier Luca Lanzi, Wolfgang Stolzmann, and Stewart W. Wilson, editors. Advances in learning classifier systems: Third International Workshop, IWLCS 2000, Paris, France, September 15–16, 2000: Revised Papers, volume 1996 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42437-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1996; Q325.5 .A344 2001. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1996.
htm;   http://www.springerlink.
com/content/978-3-540-
42437-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1996.

Lanzi:2002:ALC

[LSW02] Pier Luca Lanzi, Wolfgang Stolzmann, and Stewart W. Wilson, editors. Advances in learning classifier systems: 4th International Workshop, IWLCS 2001, San Francisco, CA, USA, July 2001: Revised Papers, volume 2321 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43793-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .A344 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2321.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43793-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2321.

Lanzi:2003:LCS

[LSW03] Pier Luca Lanzi, Wolfgang Stolzmann, and Stewart W. Wilson, editors. Learning Classifier Systems: 5th International Workshop, IWLCS 2002, Granada, Spain, September 2002: Revised Papers, volume 2661 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20544-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .I87 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2661.htm;
 http://www.springerlink.
com/content/978-3-540-
20544-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2661;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94229.

Lucks:2008:RCS

[LSW08] Stefan Lucks, Ahmad-Reza Sadeghi, and Christopher Wolf, editors. Research in Cryptology: Second Western European Workshop, WE-WoRC 2007, Bochum, Germany, July 4–6, 2007, Revised Selected Papers, volume 4945 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88352-5 (print), 3-540-88353-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88353-1.

Luo:2010:AWB

[LSW^{+}10] Xiangfeng Luo, Marc Spaniol, Lizhe Wang, Qing Li, Wolfgang Nejdl, and Wu Zhang, editors. Advances in Web-Based Learning — ICWL 2010: 9th International Conference, Shanghai, China, December 8–10, 2010. Proceedings, volume 6483 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17406-X (print), 3-642-17407-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17407-0.

Lopez:2012:CIP

[LSW12] Javier Lopez, Roberto Setola, and Stephen D. Wolthusen, editors. Critical Infrastructure Protection: Information Infrastructure Models, Analysis, and Defense, volume 7130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28919-3 (print), 3-642-28920-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28920-0.

Levi:2006:CIS

[LSY^{+}06] Albert Levi, Erkay Savaş, Hüsnü Yenigün, Selim Balcısoy, and Yücel Saygın, editors. Computer and Information Sciences — ISCIS 2006: 21th International Symposium, Istanbul, Turkey, November 1–3, 2006. Proceedings, volume 4263 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47242-8 (print), 3-540-47243-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47243-8.

Li:2008:CMI

[LSYO08] Qing Li, Stefano Spaccapietra, Eric Yu, and Antoni Olivé, editors. Conceptual Modeling — ER 2008: 27th International Conference on Conceptual Modeling, Barcelona, Spain, October 20–24, 2008. Proceedings, volume 5231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87876-9 (print), 3-540-87877-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87877-3.

Lee:2000:ACI

[LT00] D. T. Lee and Shang-Hua Teng, editors. Algorithms and Computation: 11th International Conference, ISAAC 2000, Taipei, Taiwan, December 18–20, 2000. Proceedings, volume 1969 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41255-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1969. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1969.htm;
 http://www.springerlink.
com/content/978-3-540-
41255-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1969.

LeBoudec:2001:NCT

[LT01] Jean-Yves Le Boudec and Patrick Thiran. Network calculus: a theory of deterministic queuing systems for the Internet, volume 2050 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42184-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xix + 274 pp. LCCN QA267.A1 L43 no.2050; TK5105.875.I57 L455 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2050.htm;
http://www.springerlink.
com/content/978-3-540-
42184-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2050.

Leite:2005:CLM

[LT05] João Leite and Paolo Torroni, editors. Computational Logic in Multi-Agent Systems: 5th International Workshop, CLIMA V, Lisbon, Portugal, September 29– 30, 2004, Revised Selected and Invited Papers, volume 3487 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28060-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.L63 C67 2004b.

URL http://www.springerlink.
com/content/978-3-540-
28060-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3487.

Lecomte:2011:LDI

[LT11a] Alain Lecomte and Samuel Tronçon, editors. Ludics, Dialogue and Interaction: PRELUDE Project — 2006-2009. Revised Selected Papers, volume 6505 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19210-6 (print), 3-642-19211-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19211-1.

Lopez:2011:ACN

[LT11b] Javier Lopez and Gene Tsudik, editors. Applied Cryptography and Network Security: 9th International Conference, ACNS 2011, Nerja, Spain, June 7–10, 2011. Proceedings, volume 6715 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21553-X (print), 3-642-21554-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21554-4.

Leite:2011:CLM

[LTÅ^{+}11] João Leite, Paolo Torroni, Thomas Ågotnes, Guido Boella, and Leon van der Torre, editors. Computational Logic in Multi-Agent Systems: 12th International Workshop, CLIMA XII, Barcelona, Spain, July 17–18, 2011. Proceedings, volume 6814 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22358-3 (print), 3-642-22359-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22359-4.

Lee:2011:AMMa

[LTL^{+}11a] Kuo-Tien Lee, Wen-Hsiang Tsai, Hong-Yuan Mark Liao, Tsuhan Chen, Jun-Wei Hsieh, and Chien-Cheng Tseng, editors. Advances in Multimedia Modeling: 17th International Multimedia Modeling Conference, MMM 2011, Taipei, Taiwan, January 5–7, 2011, Proceedings, Part I, volume 6523 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17831-6 (print), 3-642-17832-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17832-0.

Lee:2011:AMMb

[LTL^{+}11b] Kuo-Tien Lee, Wen-Hsiang Tsai, Hong-Yuan Mark Liao, Tsuhan Chen, Jun-Wei Hsieh, and Chien-Cheng Tseng, editors. Advances in Multimedia Modeling: 17th International Multimedia Modeling Conference, MMM 2011, Taipei, Taiwan, January 5–7, 2011, Proceedings, Part II, volume 6524 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17828-6 (print), 3-642-17829-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17829-0.

Lukowicz:2007:ACS

[LTT07] Paul Lukowicz, Lothar Thiele, and Gerhard Tröster, editors. Architecture of Computing Systems — ARCS 2007: 20th International Conference, Zurich, Switzerland, March 12–15, 2007. Proceedings, volume 4415 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71267-4 (print), 3-540-71270-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71270-1.

Lee:2006:DSA

[LTW06] Mong Li Lee, Kian Lee Tan, and Vilas Wuwongse, editors. Database Systems for Advanced Applications: 11th International Conference, DASFAA 2006, Singapore, April 12–15, 2006. Proceedings, volume 3882 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33337-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33337-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3882.

Lin:2011:CNS

[LTW11] Dongdai Lin, Gene Tsudik, and Xiaoyun Wang, editors. Cryptology and Network Security: 10th International Conference, CANS 2011, Sanya, China, December 10–12. Proceedings, volume 7092 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25512-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25512-0.

Lubich:1995:TCF

[Lub95] Hannes P. Lubich. Towards a CSCW framework for scientific cooperation in Europe, volume 889 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-58844-2 (Berlin), 0-387-58844-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 268 pp. LCCN HD66 .L83 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0889.htm;
 http://www.springerlink.
com/content/978-3-540-
58844-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=889.

Luckham:1987:ALA

[Luc87] David C. Luckham.

ANNA, a language for annotating Ada programs: reference manual, volume 260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17980-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). v + 143 pp. LCCN QA76.73.A54 A561 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0260.htm;
 http://www.springerlink.
com/content/978-0-387-
17980-3;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=260.

Spine title: ANNA refer-

ence manual.

Lukose:1997:CSF

[Luk97] Dickson Lukose, editor. Conceptual structures: fulfilling Peirce’s dream: Fifth International Conference on Conceptual Structures, ICCS ’97, Seattle, Washington, USA, August 3– 8, 1997: proceedings, volume 1257 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63308-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2.I58 1997.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1257.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63308-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1257.

Luo:2004:CDV

[Luo04] Yuhua Luo, editor. Cooperative Design, Visualization, and Engineering: First International Conference, CDVE 2004, Palma de Mallorca, Spain, September 19–22, 2004, Proceedings, volume 3190 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23149-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3190.htm;
 http://www.springerlink.
com/content/978-3-540-
23149-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3190;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100089.

Luo:2005:CDV

[Luo05] Yuhua Luo, editor. Cooperative design, visualization, and engineering: second international conference, CDVE 2005, Palma de Mallorca, Spain, September 18–21, 2005: proceedings, volume 3675 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28948-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28948-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3675.

Luo:2006:CDV

[Luo06] Yuhua Luo, editor. Cooperative Design, Visualization, and Engineering: Third International Conference, CDVE 2006, Mallorca, Spain, September 17–20, 2006. Proceedings, volume 4101 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44494-7 (print), 3-540-44496-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        44496-1.

Luo:2007:CDV

[Luo07] Yuhua Luo, editor. Cooperative Design, Visualization, and Engineering: 4th International Conference, CDVE 2007, Shanghai, China, September 16–20, 2007. Proceedings, volume 4674 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74779-6 (print), 3-540-74780-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74780-2.

Luo:2008:CDV

[Luo08] Yuhua Luo, editor. Cooperative Design, Visualization, and Engineering: 5th International Conference, CDVE 2008 Calvià, Mallorca, Spain, September 21–25, 2008 Proceedings, volume 5220 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88010-0 (print), 3-540-88011-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88011-0.

Luo:2009:CDV

[Luo09] Yuhua Luo, editor. Cooperative Design, Visualization, and Engineering: 6th International Conference, CDVE 2009, Luxembourg, Luxembourg, September 20–23, 2009. Proceedings, volume 5738 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04264-3 (print), 3-642-04265-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04265-2.

Luo:2010:CDV

[Luo10] Yuhua Luo, editor. Cooperative Design, Visualization, and Engineering: 7th International Conference, CDVE 2010, Calvia, Mallorca, Spain, September 19–22, 2010. Proceedings, volume 6240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16065-4 (print), 3-642-16066-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16066-0.

Luo:2011:CDV

[Luo11] Yuhua Luo, editor. Cooperative Design, Visualization, and Engineering: 8th International Conference, CDVE 2011, Hong Kong, China, September 11–14, 2011. Proceedings, volume 6874 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23733-9 (print), 3-642-23734-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23734-8.

Luo:2012:CDV

[Luo12] Yuhua Luo, editor. Cooperative Design, Visualization, and Engineering: 9th International Conference, CDVE 2012, Osaka, Japan, September 2–5, 2012. Proceedings, volume 7467 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32608-0 (print), 3-642-32609-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32609-7.

Lee:2002:PRS

[LV02] S.-W. Lee and Alessandro Verri, editors. Pattern Recognition with Support Vector Machines: First International Workshop, SVM 2002, Niagara Falls, Canada, August 10, 2002. Proceedings, volume 2388 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44016-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 S86 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2388.htm;
 http://www.springerlink.
com/content/978-3-540-
44016-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2388.

Li:2005:WWG

[LV05] Ki-Joune Li and Christelle Vangenot, editors. Web and wireless geographical information systems: 5th international workshop, W2GIS 2005, Lausanne, Switzerland, December 15–16, 2005: proceedings, volume 3833 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30848-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.212 .W24 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30848-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3833.

Lewenstein:2006:CPM

[LV06] Moshe Lewenstein and Gabriel Valiente, editors. Combinatorial Pattern Matching: 17th Annual Symposium, CPM 2006, Barcelona, Spain, July 5–7, 2006. Proceedings, volume 4009 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35455-7 (print), 3-540-35461-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35461-1.

Langendoen:2007:WSN

[LV07a] Koen Langendoen and Thiemo Voigt, editors. Wireless Sensor Networks: 4th European Conference, EWSN 2007, Delft, The Netherlands, January 29–31, 2007. Proceedings, volume 4373 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69829-9 (print), 3-540-69830-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69830-2.

Lumpe:2007:SCI

[LV07b] Markus Lumpe and Wim Vanderperren, editors. Software Composition: 6th International Symposium, SC 2007, Braga, Portugal, March 24–25, 2007, Revised Selected Papers, volume 4829 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77350-9 (print), 3-540-77351-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77351-1.

Lacroix:2012:RDT

[LV12] Zoé Lacroix and María Esther Vidal, editors. Resource Discovery: Third International Workshop, RED 2010, Paris, France, November 5, 2010, Revised Selected Papers, volume 6799 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27391-2 (print), 3-642-27392-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27392-6.

Lester:2004:ITS

[LVP04] James C. Lester, Rosa Maria Vicari, and Fábio Paraguaçu, editors. Intelligent Tutoring Systems: 7th International Conference, ITS 2004, Maceio, Alagoas, Brazil, August 30–September 3, 2004, Proceedings, volume 3220 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22948-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3220.htm;
 http://www.springerlink.
com/content/978-3-540-
22948-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3220;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100137.

Lammel:2008:GTT

[LVS08] Ralf Lämmel, Joost Visser, and João Saraiva, editors. Generative and Transformational Techniques in Software Engineering II: International Summer School, GTTSE 2007, Braga, Portugal, July 2–7, 2007. Revised Papers, volume 5235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88642-7 (print), 3-540-88643-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88643-3.

Li:2005:NAA

[LVW05] Zhilin Li, Lubin Vulkov, and Jerzy Waśniewski, editors. Numerical analysis and its applications: third international conference, NAA 2004, Rousse, Bulgaria, June 29–July 3, 2004. Revised selected papers, volume 3401 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24937-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA297 .N8415.

URL http://www.springerlink.
com/content/978-3-540-
24937-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3401;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106395.

Lavrac:1995:MLE

[LW95] Nada Lavrac and Stefan Wrobel, editors. Machine learning: ECML-95: 8th European Conference on Machine Learning, Heraclion, Crete, Greece, April 25–27, 1995: proceedings, volume 912 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59286-5 (Berlin), 0-387-59286-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0912.htm;
 http://www.springerlink.
com/content/978-3-540-
59286-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=912.

Leuschel:2009:IFM

[LW09] Michael Leuschel and Heike Wehrheim, editors. Integrated Formal Methods: 7th International Conference, IFM 2009, Düsseldorf, Germany, February 16–19, 2009. Proceedings, volume 5423 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00254-4 (print), 3-642-00255-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00255-7.

Lee:2011:ACA

[LW11] Dong Hoon Lee and Xiaoyun Wang, editors. Advances in Cryptology — ASIACRYPT 2011: 17th International Conference on the Theory and Application of Cryptology and Information Security, Seoul, South Korea, December 4–8. Proceedings, volume 7073 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25384-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25384-3.

Liu:2012:FHI

[LW12] Zhiming Liu and Alan Wassyng, editors. Foundations of Health Informatics Engineering and Systems: First International Symposium, FHIES 2011, Johannesburg, South Africa, August 29–30, 2011. Revised Selected Papers, volume 7151 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32354-5 (print), 3-642-32355-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32355-3.

Leung:2000:IDE

[LwCM00] Kwong Sak Leung, Lai wan Chan, and Helen Meng, editors. Intelligent Data Engineering and Automated Learning — IDEAL 2000: Data Mining, Financial Engineering, and Intelligent Agents: Second International Conference, Shatin, NT, Hong Kong, December 13–15, 2000: Proceedings, volume 1983 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41450-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55864 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1983.htm;
 http://www.springerlink.
com/content/978-3-540-
41450-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1983.

Li:2005:ADM

[LWD05] Xue Li, Shuliang Wang, and Zhao Yang Dong, editors. Advanced data mining and applications: first international conference, ADMA 2005, Wuhan, China, July 22–24, 2005: proceedings, volume 3584 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27894-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 A332 2005.

URL http://www.springerlink.
com/content/978-3-540-
27894-8;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3584.

Lei:2012:AIC

[LWDM12] Jingsheng Lei, Fu Lee Wang, Hepu Deng, and Duoqian Miao, editors. Artificial Intelligence and Computational Intelligence: 4th International Conference, AICI 2012, Chengdu, China, October 26–28, 2012. Proceedings, volume 7530 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33477-6 (print), 3-642-33478-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33478-8.

Li:2004:AWA

[LWF04] Qing Li, Guoren Wang, and Ling Feng, editors. Advances in Web-Age Information Management: 5th International Conference, WAIM 2004, Dalian, China, July 15– 17, 2004: Proceedings, volume 3129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22418-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 W325 2004.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3129.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        22418-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3129;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98703.

Leung:2008:ABL

[LWM^{+}08] Elvis Wai Chung Leung, Fu Lee Wang, Langfang Miao, Jianmin Zhao, and Jifeng He, editors. Advances in Blended Learning: Second Workshop on Blended Learning, WBL 2008, Jinhua, China, Augustl 20–22, 2008. Revised Selected Papers, volume 5328 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89961-8 (print), 3-540-89962-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89962-4.

Loog:2011:PRB

[LWRdR11] Marco Loog, Lodewyk Wessels, Marcel J. T. Reinders, and Dick de Ridder, editors. Pattern Recognition in Bioinformatics: 6th IAPR International Conference, PRIB 2011, Delft, The Netherlands, November 2–4, 2011. Proceedings, volume 7036 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24854-3 (print), 3-642-24855-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24855-9.

Liu:2009:AMT

[LWYN09] Jiming Liu, Jinglong Wu, Yiyu Yao, and Toyoaki Nishida, editors. Active Media Technology: 5th International Conference, AMT 2009, Beijing, China, October 22– 24, 2009. Proceedings, volume 5820 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04874-9 (print), 3-642-04875-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04875-3.

Lerner:2009:ALC

[LWZ09] Jürgen Lerner, Dorothea Wagner, and Katharina A. Zweig, editors. Algorithmics of Large and Complex Networks: Design, Analysis, and Simulation, volume 5515 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02093-3 (print), 3-642-02094-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02094-0.

Lin:2012:AMI

[LXH^{+}12] Weisi Lin, Dong Xu, Anthony Ho, Jianxin Wu, Ying He, Jianfei Cai, Mohan Kankanhalli, and Ming-Ting Sun, editors. Advances in Multimedia Information Processing — PCM 2012: 13th Pacific-Rim Conference on Multimedia, Singapore, December 4–6, 2012. Proceedings, volume 7674 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34777-0 (print), 3-642-34778-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34778-8.

Liu:2001:CAM

[LY01] Jiming Liu and Yiming Ye. E-Commerce Agents: Marketplace Solutions, Security Issues, and Supply and Demand, volume 2033 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41934-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 346 pp. LCCN QA267.A1 L43 no.2033; HF5548.32 .E182 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2033.htm;
 http://www.springerlink.
com/content/978-3-540-
41934-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2033.

Lakhnech:2004:FTM

[LY04] Y. (Yassine) Lakhnech and Sergio Yovine, editors. Formal techniques, modelling and analysis of timed and faulttolerant systems: joint international conferences on Formal Modelling and Analysis of Timed Systems, FORMATS 2004 and Formal Techniques in Real-Time and Fault-Tolerant Systems, FTRTFT 2004, Grenoble, France, September 22–24, 2004: proceedings, volume 3253 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23167-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65 F676 2004.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        23167-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3253;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100824.

Lim:2005:ISA

[LY05] Chae Hoon Lim and Moti Yung, editors. Information security applications: 5th international workshop, WISA 2004, Jeju Island, Korea, August 23–25, 2004: Revised selected papers, volume 3325 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24015-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 W577 2004.

URL http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3325;
 http://www.springerlink.
com/content/978-3-540-
24015-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3325;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b103174.

Lee:2012:ISA

[LY12] Dong Hoon Lee and Moti Yung, editors. Information Security Applications: 13th International Workshop, WISA 2012, Jeju Island, Korea, August 16–18, 2012, Revised Selected Papers, volume 7690 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35415-7 (print), 3-642-35416-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        35416-8.

Lee:2010:DXT

[LYBU10] Mong Li Lee, Jeffrey Xu Yu, Zohra Bellahsène, and Rainer Unland, editors. Database and XML Technologies: 7th International XML Database Symposium, XSym 2010, Singapore, September 17, 2010. Proceedings, volume 6309 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15683-5 (print), 3-642-15684-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15684-7.

Lio:2008:BIC

[LYCV08] Pietro Liò, Eiko Yoneki, Jon Crowcroft, and Dinesh C. Verma, editors. Bio-Inspired Computing and Communication: First Workshop on Bio-Inspired Design of Networks, BIOWIRE 2007 Cambridge, UK, April 2– 5, 2007 Revised Selected Papers, volume 5151 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92190-7 (print), 3-540-92191-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92191-2.

Liu:2001:AMT

[LYhL^{+}01] Jiming Liu, Pong C.

Yuen, Chun hung Li, Joseph Ng, and Toru Ishida, editors. Active Media Technology: 6th International Computer Science Conference, AMT 2001, Hong Kong, China, December 18–20, 2001. Proceedings, volume 2252 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-43035-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I578 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2252.
htm;   http://www.springerlink.
com/content/978-3-540-
43035-3;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2252.

Lipmaa:2006:ISC

[LYL06] Helger Lipmaa, Moti Yung, and Dongdai Lin, editors. Information Security and Cryptology: Second SKLOIS Conference, Inscrypt 2006, Beijing, China, November 29 — December 1, 2006. Proceedings, volume 4318 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49608-4 (print), 3-540-49610-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49610-6.

Lai:2011:ISC

[LYL11] Xuejia Lai, Moti Yung, and Dongdai Lin, editors. Information Security and Cryptology: 6th International Conference, Inscrypt 2010, Shanghai, China, October 20–24, 2010, Revised Selected Papers, volume 6584 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21517-3 (print), 3-642-21518-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21518-6.

Lee:2005:IRT

[LYMM05] Gary Geunbae Lee, Akio Yamada, Helen Meng, and Sung Hyon Myaeng, editors. Information retrieval technology: Second Asia Information Retrieval Symposium, AIRS 2005, Jeju Island, Korea, October 13–15, 2005: proceedings, volume 3689 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29186-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z699.A1 A75 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29186-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3689.

Li:2006:DMB

[LYT06] Jinyan Li, Qiang Yang, and Ah-Hwee Tan, editors. Data Mining for Biomedical Applications: PAKDD 2006 Workshop, BioDM 2006, Singapore, April 9, 2006. Proceedings, volume 3916 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33104-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33104-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3916.

Lee:2007:ISA

[LYY07] Jae Kwang Lee, Okyeon Yi, and Moti Yung, editors. Information Security Applications: 7th International Workshop, WISA 2006, Jeju Island, Korea, August 28–30, 2006, Revised Selected Papers, volume 4298 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71092-2 (print), 3-540-71093-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71093-6.

Lu:2000:WAI

[LZ00] Hongjun Lu and Aoying Zhou, editors. Webage information management: First International Conference, WAIM 2000, Shanghai, China, June 21– 23, 2000: Proceedings, volume 1846 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67627-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 W325 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1846.htm;
 http://www.springerlink.
com/content/978-3-540-
67627-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1846.

Lu:2002:AGC

[LZ02] Ruqian Lu and Songmao Zhang. Automatic Generation of Computer Animation: using AI for movie animation, volume 2160 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43114-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 380 pp. LCCN TR897.7 .L8 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2160.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43114-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2160.

Lu:2005:NMC

[LZ05] Xicheng Lu and Wei Zhao, editors. Networking and Mobile Computing: Third International Conference, ICC-NMC 2005, Zhangjiajie, China, August 2–4, 2005: Proceedings, volume 3619 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28102-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I5713 2005. URL

                                        http://springerlink.
                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3619;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        28102-3.

Lea:2008:CML

[LZ08] Doug Lea and Gianluigi Zavattaro, editors. Coordination Models and Languages: 10th International Conference, COORDINATION 2008, Oslo, Norway, June 4– 6, 2008. Proceedings, volume 5052 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68264-3 (print), 3-540-68265-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68265-3.

Li:2006:ADM

[LZhL06] Xue Li, Osmar R.

Zaïane, and Zhan huai Li, editors. Advanced Data Mining and Applications: Second International Conference, ADMA 2006, Xi’an, China, August 14–16, 2006 Proceedings, volume 4093 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37025-0 (print), 3-540-37026-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37026-0.

Lu:2011:NIPa

[LZK11a] Bao-Liang Lu, Liqing Zhang, and James Kwok, editors. Neural Information Processing: 18th International Conference, ICONIP 2011, Shanghai, China, November 13–17, 2011, Proceedings, Part I, volume 7062 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24954-X (print), 3-642-24955-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24955-6.

Lu:2011:NIPb

[LZK11b] Bao-Liang Lu, Liqing Zhang, and James Kwok, editors. Neural Information Processing: 18th International Conference, ICONIP 2011, Shanghai, China, November 13–17, 2011, Proceedings, Part II, volume 7063 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24957-4 (print), 3-642-24958-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24958-7.

Lu:2011:NIPc

[LZK11c] Bao-Liang Lu, Liqing Zhang, and James Kwok, editors. Neural Information Processing: 18th International Conference, ICONIP 2011, Shanghai, China, November 13–17, 2011, Proceedings, Part III, volume 7064 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24964-7 (print), 3-642-24965-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24965-5.

Lai:2011:ISI

[LZL11] Xuejia Lai, Jianying Zhou, and Hui Li, editors. Information Security: 14th International Conference, ISC 2011, Xi’an, China, October 26–29, 2011. Proceedings, volume 7001 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24860-8 (print), 3-642-24861-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24861-0.

Liu:2011:ANNa

[LZP^{+}11a] Derong Liu, Huaguang Zhang, Marios Polycarpou, Cesare Alippi, and Haibo He, editors. Advances in Neural Networks — ISNN 2011: 8th International Symposium on Neural Networks, ISNN 2011, Guilin, China, May 29– June 1, 2011, Proceedings, Part I, volume 6675 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21104-6 (print), 3-642-21105-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21105-8.

Liu:2011:ANNb

[LZP^{+}11b] Derong Liu, Huaguang Zhang, Marios Polycarpou, Cesare Alippi, and Haibo He, editors. Advances in Neural Networks — ISNN 2011: 8th International Symposium on Neural Networks, ISNN 2011, Guilin, China, May 29–June 1, 2011, Proceedings, Part II, volume 6676 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21089-9 (print), 3-642-21090-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21090-7.

Liu:2011:ANNc

[LZP^{+}11c] Derong Liu, Huaguang Zhang, Marios Polycarpou, Cesare Alippi, and Haibo He, editors. Advances in Neural Networks — ISNN 2011: 8th International Symposium on Neural Networks, ISNN 2011, Guilin, China, May 29–June 1, 2011, Proceedings, Part III, volume 6677 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21110-0 (print), 3-642-21111-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21111-9.

Li:2008:AWB

[LZS^{+}08] Frederick Li, Jianmin Zhao, Timothy K. Shih, Rynson Lau, Qing Li, and Dennis McLeod, editors. Advances in Web Based Learning — ICWL 2008: 7th International Conference, Jinhua, China, August 20–22, 2008. Proceedings, volume 5145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85032-5 (print), 3-540-85033-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85033-5.

Min:2006:FHP

[M^{+}06] Geyong Min et al., editors. Frontiers of high performance computing and networking: ISPA 2006 workshops: ISPA 2006 international workshops, FHPCN, XHPC, S-GRACE, GridGIS, HPC-GTP, PDCE, ParDM-Com, WOMP, ISDF, and UPWN, Sorrento, Italy, December 4–7, 2006: proceedings, volume 4331 of Lecture notes in computer science, 0302-9743. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-49860-5 (paperback). LCCN QA76.58 .I88 2006a; QA76.88 .F76 2006. URL http:/

/www.loc.gov/catdir/
enhancements/fy0825/
2006937143-d.html;
http://www.springerlink.
com/content/978-3-540-
49860-5.

Marshall:2003:MMN

[MA03a] Alan Marshall and Nazim Algoulmine, editors. Management of Multimedia Networks and Services: 6th IFIP/IEEE International Conference, MMNS 2003, Belfast, Northern Ireland, UK, September 7–10, 2003: Proceedings, volume 2839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20050-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.15. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2839.htm;
 http://www.springerlink.
com/content/978-3-540-
20050-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2839;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b13721.

Mira:2003:IWCa

[MA03b] José Mira and José R. (José Ramón) Alvarez, editors. 7th International Work Conference on Artificial and Natural Neural Networks: IWANN 2003, Maó, Minorca, Spain, June 3–6, 2003. Part 1, Computational Methods in Neural Modelling, volume 2686 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40210-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP363.3 .I586 2003 pt. 1–2 (2003). URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2686.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40210-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2686.

Mira:2003:IWCb

[MA03c] José Mira and José R. (José Ramón) Alvarez, editors. 7th International Work Conference on Artificial and Natural Neural Networks: IWANN 2003, Maó, Minorca, Spain, June 3–6, 2003.

Part 2. Artificial Neural Nets Problem Solving Methods, volume 2687 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40211-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP363.3 .I586 2003 pt. 1–2 (2003). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2687.
htm;   http://www.springerlink.
com/content/978-3-540-
40211-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2687.

Mira:2005:AIK

[MÁ05a] José Mira and José R.

Álvarez, editors. Artificial Intelligence and Knowledge Engineering Applications: A Bioinspired Approach: First International Work-Conference on the Interplay Between Natural and Artificial Computation, IWINAC 2005, Las Palmas, Canary Islands, Spain, June 15–18, 2005, Proceedings, Part II, volume 3562 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26319-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I584 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26319-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3562.

Mira:2005:MSM

[MÁ05b] José Mira and José R. Álvarez, editors. Mechanisms, Symbols, and Models Underlying Cognition: First International Work-Conference on the Interplay Between Natural and Artificial Computation, IWINAC 2005, Las Palmas, Canary Islands, Spain, June 15–18, 2005, Proceedings, Part I, volume 3561 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26298-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP363.3 .I587 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26298-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3561.

Mobasher:2005:ITW

[MA05c] Bamshad Mobasher and Sarabjot Singh Anand, editors. Intelligent Techniques for Web Personalization: IJCAI 2003 Workshop, ITWP 2003, Acapulco, Mexico, August 11, 2003, Revised Selected Papers, volume 3169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29846-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29846-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3169.

Mira:2007:BIM

[MÁ07a] José Mira and José R.

Álvarez, editors. Bioinspired Modeling of Cognitive Tasks: Second International Work-Conference on the Interplay Between Natural and Artificial Computation, IWINAC 2007, La Manga del Mar Menor, Spain, June 18–21, 2007, Proceedings, Part I, volume 4527 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73052-4 (print), 3-540-73053-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73053-8.

Mira:2007:NIP

[MÁ07b] José Mira and José R. Álvarez, editors. Nature Inspired Problem-Solving Methods in Knowledge Engineering: Second International Work-Conference on the Interplay Between Natural and Artificial Computation, IWINAC 2007, La Manga del Mar Menor, Spain, June 18– 21, 2007, Proceedings, Part II, volume 4528 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73054-0 (print), 3-540-73055-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73055-2.

Munch:2007:PFS

[MA07c] Jürgen Münch and Pekka Abrahamsson, editors. Product-Focused Software Process Improvement: 8th International Conference, PROFES 2007, Riga, Latvia, July 2–4, 2007. Proceedings, volume 4589 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73459-7 (print), 3-540-73460-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73460-4.

Metaxas:2011:FIM

[MA11] Dimitris N. Metaxas and Leon Axel, editors. Functional Imaging and Modeling of the Heart: 6th International Conference, FIMH 2011, New York City, NY, USA, May 25–27, 2011. Proceedings, volume 6666 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21027-9 (print), 3-642-21028-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21028-0.

Monien:1991:ALP

[MAA91] B. Monien, J. Leach Albert, and M. Rodriquez Artalejo, editors. Automata, languages and programming: 18th international colloquium, Madrid, Spain, July 8–12, 1991: proceedings, volume 510 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54233-7 (Berlin), 0-387-54233-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1.L43 no.510. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0510.htm;
 http://www.springerlink.
com/content/978-3-540-
54233-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=510.

ICALP ’91 (International Colloquium on Automata, Languages, and Programming) was sponsored by the European Association for Theoretical Computer Science (EATCS) and Universidad Complutense de Madrid.

MereloGuervos:2002:PPS

[MAB^{+}02] Juan Julián Merelo Guervós, Panagiotis Adamidis, Hans-Georg Beyer, Josè-Luis Fernández-Villacañas, and Hans-Paul Schwefel, editors. Parallel problem solving from nature

— PPSN VII: 7th International Conference, Granada, Spain, September 7–11, 2002: Proceedings, volume 2439 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44139-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I535 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2439.htm;
 http://www.springerlink.
com/content/978-3-540-
44139-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2439.

MacLean:2006:SCV

[Mac06] W. James MacLean, editor. Spatial coherence for visual motion analysis: first international workshop, SCVMA 2004, Prague, Czech Republic, May 15, 2004. revised papers, volume 3667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32533-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32533-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3667.

MacKinnon:2012:DSS

[Mac12] Lachlan M. MacKinnon, editor. Data Security and Security Data: 27th British National Conference on Databases, BN-COD 27, Dundee, UK, June 29 — July 1, 2010. Revised Selected Papers, volume 6121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25703-8 (print), 3-642-25704-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25704-9.

Madhavan:1989:FST

[Mad89] C. E. Veni Madhavan, editor. Foundations of software technology and theoretical computer science: ninth conference, Bangalore, India, December 19–21, 1989: proceedings, volume 405 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-52048-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F681 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0405.htm;
 http://www.springerlink.
com/content/978-0-387-
52048-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=405.

Proceedings of the ninth annual FST and TCS conference. “FST and TCS 9.”.

Madsen:1992:EEC

[Mad92] O. Lehrmann Madsen, editor. ECOOP ’92, European Conference on Object-Oriented Programming: Utrecht, The Netherlands, June 29–July 3, 1992: proceedings, volume 615 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55668-0 (Berlin), 0-387-55668-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0615.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        55668-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=615.

Manandhar:2004:ACS

[MAD^{+}04] Suresh Manandhar, Jim Austin, Uday Desai, Yoshio Oyanagi, and Asoke Talukder, editors. Applied computing: Second Asian Applied Computing Conference, AACC 2004, Kathmandu, Nepal, October 29–31, 2004: proceedings, volume 3285 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23659-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.A65 A82 2004.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        23659-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3285.

Metaxas:2008:MICa

[MAFS08a] Dimitris Metaxas, Leon Axel, Gabor Fichtinger, and Gábor Székely, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2008: 11th International Conference, New York, NY, USA, September 6–10, 2008, Proceedings, Part I, volume 5241 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85987-X (print), 3-540-85988-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85988-8.

Metaxas:2008:MICb

[MAFS08b] Dimitris Metaxas, Leon Axel, Gabor Fichtinger, and Gábor Székely, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2008: 11th International Conference, New York, NY, USA, September 6–10, 2008, Proceedings, Part II, volume 5242 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85989-6 (print), 3-540-85990-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85990-1.

Monroy:2004:MAA

[MAFSS04] Raúl Monroy, Gustavo Arroyo-Figueroa, Luis Enrique Sucar, and Humberto Sossa, editors. MICAI 2004: Advances in Artificial Intelligence: Third Mexican International Conference on Artificial Intelligence, Mexico City, Mexico, April 26–30, 2004: Proceedings, volume 2972 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21459-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2972.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21459-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2972;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96521.

Magnusson:1998:SCM

[Mag98] Boris Magnusson, editor. System configuration management: ECOOP’98 SCM-8 Symposium, Brussels, Belgium, July 20–21, 1998:

Proceedings, volume 1439 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64733-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C69 E26 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1439.htm;
 http://www.springerlink.
com/content/978-3-540-
64733-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1439.

Magnusson:2002:EOO

[Mag02] Boris Magnusson, editor.

ECOOP 2002 — objectoriented programming: 16th European Conference, Malaga, Spain, June 10–14, 2002: Proceedings, volume 2374 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43759-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2374.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        43759-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2374.

Magnin:2003:FIM

[Mag03] Isabelle F. Magnin, editor. Functional imaging and modeling of the heart: Second International Workshop, FIMH 2003, Lyon, France, June 5–6, 2003: Proceedings, volume 2674 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40262-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP112.4 .I58 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2674.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40262-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2674.

Maheshwari:1985:FST

[Mah85] S. N. Maheshwari, editor. Foundations of software technology and theoretical computer science: fifth conference, New Delhi, India, December 16–18, 1985: proceedings, volume 206 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-16042-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F681 1985. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0206.htm;
 http://www.springerlink.
com/content/978-0-387-
16042-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=206.

Fifth Conference on Foundations of Software Technology and Theoretical Computer Science sponsored jointly by Indian Institute of Technology Delhi and Tata Institute of Fundamental Research, Bombay.

Maher:2004:ACS

[Mah04] Michael J. Maher, editor.

Advances in Computer Science: ASIAN 2004: Higher-Level Decision Making. 9th Asian Computing Science Conference. Dedicated to Jean-Louis Lassez on the Occasion of His 5th Cycle [60th] Birthday. Chiang Mai, Thailand, December 8–10, 2004. Proceedings, volume 3321 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24087-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24087-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3321;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b103476.

Main:1988:MFP

[Mai88] M. Main, editor. Mathematical foundations of programming language semantics: 3rd workshop, Tulane University, New Orleans, Louisiana, USA, April 8–10, 1987: proceedings, volume 298 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19020-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .M3781 1988. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0298.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        19020-4;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=298.

“The Third Workshop on the Mathematical Foundations of Programming Language Semantics … ”–Pref.

Main:1990:MFP

[Mai90] M. Main, editor. Mathematical foundations of programming semantics: 5th international conference, Tulane University, New Orleans, Louisiana, USA, March 29–April 1, 1989: proceedings, volume 442 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52790-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .M38 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0442.htm;
 http://www.springerlink.
com/content/978-0-387-
52790-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=442.

Proceedings of the Fifth International Conference on the Mathematical Foundations of Programming Semantics.

Maibaum:2000:FAS

[Mai00] Thomas S. E. Maibaum, editor. Fundamental approaches to software engineering: Third International Conference, FASE 2000 held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2000, Berlin, Germany, March/April 2000: Proceedings, volume 1783 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67261-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1783.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1783.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67261-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1783.

Makedon:1986:VAA

[Mak86] F. Makedon, editor. VLSI algorithms and architectures: Aegean Workshop on Computing, Loutraki, Greece, July 8–11, 1986: proceedings, volume 227 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16766-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874 .A3381 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0227.htm;
 http://www.springerlink.
com/content/978-0-387-
16766-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=227.

Malyshkin:1995:PCT

[Mal95] V. E. Malyshkin, editor. Parallel computing technologies: third international conference, PaCT-95, St. Petersburg, Russia, September 12–25, 1995: proceedings, volume 964 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60222-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I547 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0964.htm;
 http://www.springerlink.
com/content/978-3-540-
60222-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=964.

Maler:1997:HRT

[Mal97a] O. (Oded) Maler, editor. Hybrid and realtime systems: International Workshop, HART ’97, Grenoble, France, March 26–28, 1997: proceedings, volume 1201 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62600-X (Berlin softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38 .H37 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1201.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62600-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1201.

Malyshkin:1997:PCT

[Mal97b] V. E. (Viktor Emmanuilovich) Malyshkin, editor. Parallel computing technologies: 4th international conference, PaCT-97, Yaroslavl, Russia, September 1997: proceedings, volume 1277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63371-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I547 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1277.htm;
 http://www.springerlink.
com/content/978-3-540-
63371-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1277.

Malyshkin:1999:PCT

[Mal99] V. E. (Viktor Emmanilovich) Malyshkin, editor. Parallel computing technologies: 5th International Conference, PaCT-99, St. Petersburg, Russia, September 6–10, 1999: Proceedings, volume 1662 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66363-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I547 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1662.htm;
 http://www.springerlink.
com/content/978-3-540-
66363-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1662.

Malyshkin:2001:PCT

[Mal01] Victor E. Malyshkin, editor. Parallel computing technologies: 6th International Conference, PaCT 2001, Novosibirsk, Russia, September 3–7, 2001: Proceedings, volume 2127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42522-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2127; QA76.58 .I547 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2127.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42522-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2127.

Malkhi:2002:DCI

[Mal02] Dahlia Malkhi, editor. Distributed Computing: 16th International Conference, DISC 2002, Toulouse, France, October 28–30, 2002. Proceedings, volume 2508 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00073-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D438 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2508.htm;
 http://www.springerlink.
com/content/978-3-540-
00073-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2508.

Also available via the World Wide Web.

Malyshkin:2003:PCT

[Mal03] Viktor (Viktor Emmanuilovich) Malyshkin, editor. Parallel Computing Technologies: 7th International Conference, PaCT 2003, Nizhni Novgorod, Russia, September 15–19, 2003: Proceedings, volume 2763 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40673-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2763.htm;
 http://www.springerlink.
com/content/978-3-540-
40673-0;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2763;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11928.

Malyshkin:2005:PCT

[Mal05] Victor Malyshkin, editor. Parallel computing technologies: 8th international conference, PaCT 2005, Krasnoyarsk, Russia, September 5–9, 2005: proceedings, volume 3606 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28126-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I547 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28126-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3606.

Malyshkin:2007:PCT

[Mal07] Victor Malyshkin, editor. Parallel Computing Technologies: 9th International Conference, PaCT 2007, Pereslavl-Zalessky, Russia, September 3–7, 2007. Proceedings, volume 4671 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73939-4 (print), 3-540-73940-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73940-1.

Malkin:2008:TCC

[Mal08] Tal Malkin, editor. Topics in Cryptology — CT-RSA 2008: The Cryptographers’ Track at the RSA Conference 2008, San Francisco, CA, USA, April 8–11, 2008. Proceedings, volume 4964 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79262-7 (print), 3-540-79263-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79263-5.

Malyshkin:2009:PCT

[Mal09] Victor Malyshkin, editor. Parallel Computing Technologies: 10th International Conference, PaCT 2009, Novosibirsk, Russia, August 31– September 4, 2009. Proceedings, volume 5698 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03274-5 (print), 3-642-03275-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03275-2.

Malyshkin:2011:PCT

[Mal11] Victor Malyshkin, editor. Parallel Computing Technologies: 11th International Conference, PaCT 2011, Kazan, Russia, September 19– 23, 2011. Proceedings, volume 6873 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23177-2 (print), 3-642-23178-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23178-0.

Manes:1975:CTA

[Man75] Ernest G. Manes, editor. Category theory applied to computation and control: proceedings of the First International Symposium, San Francisco, February 25–26, 1974, volume 25 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07142-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .C351. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0025.htm;
 http://www.springerlink.
com/content/978-0-387-
07142-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=25.

“An A.M.S. symposium within the annual meeting of the American Association for the Advancement of Science.”.

Maneth:2009:IAA

[Man09] Sebastian Maneth, editor. Implementation and Application of Automata: 14th International Conference, CIAA 2009, Sydney, Australia, July 14–17, 2009. Proceedings, volume 5642 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02978-7 (print), 3-642-02979-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02979-0.

Marchuk:1975:OTP

[Mar75] G. I. Marchuk, editor. Optimization techniques: [proceedings], volume 27 of Series, I.F.I.P. TC 7 optimization conferences and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07165-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA402.5 .I181 1974. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0027.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        07165-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=27.

Marlin:1980:CPM

[Mar80] Christopher D. Marlin. Coroutines: a programming methodology, a language design, and an implementation, volume 95 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10256-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 246 pp. LCCN QA76.6 .M3595 1980. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0095.htm;
 http://www.springerlink.
com/content/978-0-387-
10256-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=95.

Originally presented as the author’s thesis, University of Adelaide.

Marsan:1993:ATP

[Mar93] M. Ajmone Marsan, editor. Application and theory of Petri nets, 1993: 14th International Conference, Chicago, Illinois, USA, June 21– 25, 1993: proceedings, volume 691 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56863-8 (Berlin), 0-387-56863-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I574 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0691.htm;
 http://www.springerlink.
com/content/978-3-540-
56863-6;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=691.

“Proceedings of the 14th International Conference on Application and Theory of Petri Nets”–Pref.

Marks:2001:GDI

[Mar01] Joe Marks, editor. Graph drawing: 8th International Symposium GD 2000, Colonial Williamsburg, Va, USA, September 20–23, 2000: Proceedings, volume 1984 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41554-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 2000.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1984.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        41554-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1984.

Marsan:2003:QSM

[Mar03] M. Ajmone Marsan, editor. Quality of service in multiservice IP networks: Second International Workshop, QoS-IP 2003, Milano, Italy, February 24–26, 2003: Proceedings, volume 2601 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00604-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.75 .Q67 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2601.htm;
 http://www.springerlink.
com/content/978-3-540-
00604-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2601.

Also available via the World Wide Web.

Margenstern:2005:MCU

[Mar05a] Maurice Margenstern, editor. Machines, Computations, and Universality: 4th International Conference, MCU 2004, Saint Petersburg, Russia, September 21–24, 2004: Revised selected papers, volume 3354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25261-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .M29 2004. URL

http://springerlink.
metapress.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3354;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        25261-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3354;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106980.

Muanoz-Avila:2005:CBR

[MAR05b] Héctor Muänoz-Avila and Francesco Ricci, editors. Case-Based Reasoning Research and Development: 6th International Conference on Case-Based Reasoning, ICCBR 2005, Chicago, IL, USA, August 23– 26, 2005: Proceedings, volume 3620 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28174-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I54 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28174-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3620.

Marcus:2011:DUEa

[Mar11a] Aaron Marcus, editor.

Design, User Experience, and Usability. Theory, Methods, Tools and Practice: First International Conference, DUXU 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part I, volume 6769 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21674-9 (print), 3-642-21675-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21675-6.

Marcus:2011:DUEb

[Mar11b] Aaron Marcus, editor.

Design, User Experience, and Usability. Theory, Methods, Tools and Practice: First International Conference, DUXU 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part II, volume 6770 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21707-9 (print), 3-642-21708-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21708-1.

Marino:2011:FCL

[Mar11c] Julio Mariño, editor. Functional and Constraint Logic Programming: 19th International Workshop, WFLP 2010, Madrid, Spain, January 17, 2010. Revised Selected Papers, volume 6559 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20774-X (print), 3-642-20775-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20775-4.

Mattos:1991:AKB

[Mat91] Nelson Mendonca Mattos. An approach to knowledge base management, volume 513 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54268-X (Berlin), 0-387-54268-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 247 pp. LCCN QA76.76.E95 M3 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0513.
htm;   http://www.springerlink.
com/content/978-3-540-
54268-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=513.

Matousek:2001:TSD

[Mat01] Václav Matoušek, editor. Text, speech, and dialogue: 4th International Conference, TSD 2001, Zelezna Ruda, Czech Republic, September 11–13, 2001: Proceedings, volume 2166 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42557-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 T47 2001; QA267.A1 L43 no.2166.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2166.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        42557-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2166.

Matsui:2002:FSE

[Mat02] Mitsuru Matsui, editor. Fast software encryption: 8th International Workshop, FSE 2001, Yokohama, Japan, April 2– 4, 2001: Revised Papers, volume 2355 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43869-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F77 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2355.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43869-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2355.

Matsui:2009:ACA

[Mat09] Mitsuru Matsui, editor. Advances in Cryptology

— ASIACRYPT 2009: 15th International Conference on the Theory and Application of Cryptology and Information Security, Tokyo, Japan, December 6–10, 2009.

Proceedings, volume 5912 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10365-0 (print), 3-642-10366-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10366-7.

Maurer:1979:ALP

[Mau79] Hermann A. Maurer, editor. Automata, languages and programming: sixth colloquium, Graz, Austria, July 16– 20, 1979, volume 71 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CO-DEN LNCSD9. ISBN 0-387-09510-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.71.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0071.
htm;   http://www.springerlink.
com/content/978-0-387-
09510-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=71.

Maurer:1989:CAL

[Mau89] Hermann A. Maurer, editor. Computer assisted learning: 2nd International Conference, IC-CAL ’89, Dallas, TX, USA, May 9–11, 1989: proceedings, volume 360 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51142-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.5 .I1161 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0360.htm;
 http://www.springerlink.
com/content/978-0-387-
51142-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=360.

Maurer:1991:NRN

[Mau91] Hermann A. Maurer, editor. New results and new trends in computer science, Graz, Austria, June 20–21, 1991: proceedings, volume 555 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54869-6 (Berlin), 0-387-54869-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .N49 1991. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t0555.htm;
http://www.springerlink.
com/content/978-3-540-
54869-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=555.

Maurer:1996:ACE

[Mau96] Ueli Maurer, editor. Advances in cryptology, EUROCRYPT ’96: International Conference on the Theory and Application of Cryptographic Techniques, Saragossa, Spain, May 12–16, 1996: proceedings, volume 1070 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61186-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1070.
htm;   http://www.springerlink.
com/content/978-3-540-
61186-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1070.

Sponsored by the International Association for Cryptologic Research (IACR), in cooperation with the University of Saragossa.

Mayr:1993:GTC

[May93] Ernst Mayr, editor. Graph-theoretic concepts in computer science: 18th International Workshop, WG ’92, Wiesbaden-Naurod, Germany, June 18–20, 1992: proceedings, volume 657 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56402-0 (New York), 3-540-56402-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G715 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0657.htm;
 http://www.springerlink.
com/content/978-0-387-
56402-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=657.

Proceedings of the 18th International Workshop on Graph-Theoretic Concepts in Computer Science. Contents: Restricted Graph Classes. On improved time bounds for permutation graph problems / A. Brandstadt. A simple test for interval graphs / W.-L. Hsu. Tolerance graphs and orders / S. Felsner — Scheduling and Related Problems. On scheduling problems restricted to interval orders / K. Jansen. Scheduling with incompatible jobs / H. L. Bodlaender, K. Jansen and G. J. Woeginger. Generalized coloring for tree-like graphs / K. Jansen and P. Scheffler — Parallel and Distributed Algorithms I. Optimal (parallel) algorithms for the all-to-all vertices distance problem for certain graph classes / E. Dahlhaus. Topology of parallel networks and computational complexity / J. Hromkovic. Parallel triangulation of nonconvex polytopes / W. Preilowski — Combinatorial Graph Problems I. Kayles on special classes of graphs — An application of Sprague-Grundy theory / H. L. Bodlaender. A linear time algorithm for isomorphism of graphs of bounded average genus / J. Chen. Contents: Improved algorithms for routing on two-dimensional grids / D. Bhatia, T. Leighton, F. Makedon and C. H. Norton. Minimum rectilinear Steiner trees for intervals on two parallel lines / E. Ihler — Parallel and Distributed Algorithms II. A new characterization of tree medians with applications to distributed algorithms / O. Gerstel and S. Zaks — Graph Decomposition. The 3-edge-components and a structural description of all 3-edge-cuts in a graph / E. Dinitz. On assembly of 4-connected graphs / J. Chen and A. Kanevsky. On the homogeneous decomposition of graphs / B. Jamison and S. Olariu — Combinatorial Graph Problems II. Embeddings in recursive combinatorial networks / S. K. Das and A. Mao. On shortcutting digraphs / M. Thorup. An efficient algorithm to recognize prime undirected graphs / A. Cournier and M. Habib. On the complexity of partial order properties / S. Felsner and D. Wagner Graph Grammars and Geometry. Probabilistic graph grammars / M. Mosbah. Contents: Single vs. double pushout derivations of graphs / F. Parisi-Presicce. Hexagonal grid drawings / G. Kant Modelling by Graphs. Graph algorithms = iteration + data structures? The structure of graph algorithms and a corresponding style of programming / M. Erwig. Petri nets, hypergraphs and conflicts / P. Alimonti and E. Feuerstein. Analysis and manipulation of Boolean functions in terms of decision graphs / J. Gergov and C. Meinel. The expressiveness of silence: Tight bounds for synchronous communication of information using bits and silence / U.-M. O’Reilly and N. Santoro. The power and the limitations of local computations on graphs / I. Litovsky, Y. Metivier and W. Zielonka.

Mayr:2001:DES

[May01] H. C. Mayr, editor.

Database and expert systems applications: 12th International Conference, DEXA 2001, Munich, Gemany, September 3–5, 2001: Proceedings, volume 2113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42527-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D32 D363 2001; Full-text. Restricted to UC campuses. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2113.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        42527-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2113.

Mazurkiewicz:1976:MFC

[Maz76] Antoni W. Mazurkiewicz, editor. Mathematical foundations of computer science, 1976: proceedings, 5th Symposium, Gdansk, September 6– 10, 1976, volume 45 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1976. CO-DEN LNCSD9. ISBN 0-387-07854-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.S4 S989 1976.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0045.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        07854-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=45.

“The Symposium is being organized by the Computation Centre of the Polish Academy of Sciences in cooperation with the University of Gdansk.”.

Maekawa:1982:OSE

[MB82] M. Maekawa and L. A. Belady, editors. Operating systems engineering: proceedings of the 14th IBM Computer Science Symposium, Amagi, Japan, October 1980, volume 143 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11604-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I15 1980. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0143.htm;
 http://www.springerlink.
com/content/978-0-387-
11604-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=143.

Muller:1987:NOS

[MB87] Gunter Muller and Robert P.

Blanc, editors. Networking in open systems: international seminar, Oberlech, Austria, August 18–22, 1986: proceedings, volume 248 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17707-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .N4671 1986. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t0248.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        17707-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=248.

“Organized jointly by the IBM European Networking Center (ENC) Heidelberg and the National Bureau of Standards (NBS), Gaithersburg, USA”–Pref.

Moller:1996:LCS

[MB96] Faron Moller and G. M. (Graham M.) Birtwistle. Logics for concurrency: structure versus automata, volume 1043 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60915-6. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 266 pp. LCCN QA76.9.D5 M645 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1043.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60915-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1043.

Murshed:1997:ADI

[MB97] Nabeel A. Murshed and Flavio Bortolozzi, editors. Advances in document image analysis: First Brazilian Symposium, BSDIA’97, Curitabia, Brazil, November 2–5, 1997: proceedings, volume 1339 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63791-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1640.B73 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1339.
htm;   http://www.springerlink.
com/content/978-3-540-
63791-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1339.

Mueller:1998:LCT

[MB98] Frank Mueller and Azer Bestavros, editors. Languages, compilers, and tools for embedded systems: ACM SIGPLAN Workshop LCTES ’98, Montréal, Canada, June 9–12, 1998: Proceedings, volume 1474 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65075-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6.A2783 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1474.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65075-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1474.

Marsan:2001:QSM

[MB01] M. Ajmone Marsan and Andrea Bianco, editors. Quality of service in multiservice IP networks: International Workshop, QoS-IP 2001, Rome, Italy, January 24–26, 2001: Proceedings, volume 1989 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41512-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.75 .Q67 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1989.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41512-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1989.

McGookin:2006:HAI

[MB06] David McGookin and Stephen Brewster, editors. Haptic and Audio Interaction Design: First International Workshop, HAID 2006, Glasgow, UK, August 31 — September 1, 2006. Proceedings, volume 4129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37595-3 (print), 3-540-37596-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37596-8.

Meyer:2009:KRA

[MB09] John-Jules Ch. Meyer and Jan Broersen, editors. Knowledge Representation for Agents and Multi-Agent Systems: First International Workshop, KRA-MAS 2008, Sydney, Australia, September 17, 2008, Revised Selected Papers, volume 5605 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05300-9 (print), 3-642-05301-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05301-6.

Moro:2005:APPb

[MBA05] Gianluca Moro, Sonia Bergamaschi, and Karl Aberer, editors. Agents and peer-to-peer computing: third international workshop, AP2PC 2004, New York, NY, USA, July 19, 2004: revised and invited papers, volume 3601 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29755-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN K5105.525 .A615 2004.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        29755-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3601.

Mullery:1997:ISN

[MBC^{+}97] Al Mullery, Michel Besson, Mario Campolargo, Roberta Gobbi, and Rick Reed, editors. Intelligence in services and networks: technology for cooperative competition: Fourth International Conference on Intelligence in Services and Networks, IS&N ’97, Cernobbio, Italy, May 27–29, 1997: proceedings, volume 1238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63135-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.A1 I47985 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1238.htm;
 http://www.springerlink.
com/content/978-3-540-
63135-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1238.

Morrison:2008:SAS

[MBF08] Ron Morrison, Dharini Balasubramaniam, and Katrina Falkner, editors. Software Architecture: Second European Conference, ECSA 2008 Paphos, Cyprus, September 29–October 1, 2008 Proceedings, volume 5292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88029-1 (print), 3-540-88030-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88030-1.

Maidment:2012:BII

[MBG12] Andrew D. A. Maidment, Predrag R. Bakic, and Sara Gavenonis, editors. Breast Imaging: 11th International Workshop, IWDM 2012, Philadelphia, PA, USA, July 8–11, 2012. Proceedings, volume 7361 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31270-5 (print), 3-642-31271-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31271-7.

Markowitch:2009:IST

[MBH^{+}09] Olivier Markowitch, Angelos Bilas, Jaap-Henk Hoepman, Chris J. Mitchell, and Jean-Jacques Quisquater, editors. Information Security Theory and Practice. Smart Devices, Pervasive Systems, and Ubiquitous Networks: Third IFIP WG 11.2 International Workshop, WISTP 2009, Brussels, Belgium, September 1– 4, 2009, Proceedings, volume 5746 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03943-X (print), 3-642-03944-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03944-7.

Mutlu:2011:SRT

[MBH^{+}11] Bilge Mutlu, Christoph Bartneck, Jaap Ham, Vanessa Evers, and Takayuki Kanda, editors. Social Robotics: Third International Conference, ICSR 2011, Amsterdam, The Netherlands, November 24– 25, 2011. Proceedings, volume 7072 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25503-5 (print), 3-642-25504-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25504-5.

Moro:2007:DIS

[MBJ^{+}07] Gianluca Moro, Sonia Bergamaschi, Sam Joseph, Jean-Henry Morin, and Aris M. Ouksel, editors. Databases, Information Systems, and Peer-to-Peer Computing: International Workshops, DBISP2P 2005/2006, Trondheim, Norway, August 28– 29, 2005, Seoul, Korea, September 11, 2006, Revised Selected Papers, volume 4125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71660-2 (print), 3-540-71661-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71661-7.

Marti:2007:PRIa

[MBMS07a] Joan Martí, José Miguel Benedí, Ana Maria Mendonça, and Joan Serrat, editors. Pattern Recognition and Image Analysis: Third Iberian Conference, IbPRIA 2007, Girona, Spain, June 6–8, 2007, Proceedings, Part I, volume 4477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72846-5 (print), 3-540-72847-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72847-4.

Marti:2007:PRIb

[MBMS07b] Joan Martí, José Miguel Benedí, Ana Maria Mendonça, and Joan Serrat, editors. Pattern Recognition and Image Analysis: Third Iberian Conference, IbPRIA 2007, Girona, Spain, June 6–8, 2007, Proceedings, Part II, volume 4478 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72848-1 (print), 3-540-72849-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72849-8.

McCune:2011:TTC

[MBP^{+}11] Jonathan M. McCune, Boris Balacheff, Adrian Perrig, Ahmad-Reza Sadeghi, Angela Sasse, and Yolanta Beres, editors. Trust and Trustworthy Computing: 4th International Conference, TRUST 2011, Pittsburgh, PA, USA, June 22–24, 2011. Proceedings, volume 6740 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21598-X (print), 3-642-21599-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21599-5.

Martin:2005:MSX

[MBS05a] R. R. (Ralph R.) Martin, Helmut Bez, and M. A. (Malcolm Arthur) Sabin, editors. Mathematics of surfaces XI: 11th IMA international conference, Loughborough, UK, September 5–7, 2005: proceedings, volume 3604 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28225-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA571; QA571 .I43 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28225-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3604.

Morik:2005:LPD

[MBS05b] Katharina Morik, Jean-François Boulicaut, and Arno Siebes, editors. Local Pattern Detection: International Seminar, Dagstuhl Castle, Germany, April 12–16, 2004, Revised Selected Papers, volume 3539 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26543-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 I587 2005.

URL http://www.springerlink.
com/content/978-3-540-
26543-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3539.

Malloy:2011:SLE

[MBS11] Brian Malloy, Mark Brand, and Steffen Staab, editors. Software Language Engineering: Third International Conference, SLE 2010, Eindhoven, The Netherlands, October 12–13, 2010, Revised Selected Papers, volume 6563 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19439-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19439-9;   http://www.
                                        springerlink.com/content/
                                        978-3-642-19439-9/.

Mamede:2003:CPP

[MBT^{+}03] Nuno J. Mamede, Jorge Baptista, Isabel Trancoso, Maria das Graças, and Volpe Nunes, editors. Computational Processing of the Portuguese Language: 6th International Workshop, PROPOR 2003 Faro, Portugal, June 26–27, 2003 Proceedings, volume 2721. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40436-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN PC5044.5 .I58 2003. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        40436-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2721.

Masip-Bruin:2011:WWI

[MBVTY11] Xavier Masip-Bruin, Dominique Verchere, Vassilis Tsaoussidis, and Marcelo Yannuzzi, editors. Wired/Wireless Internet Communications: 9th IFIP TC 6 International Conference, WWIC 2011, Vilanova i la Geltrú, Spain, June 15–17, 2011. Proceedings, volume 6649 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21559-9 (print), 3-642-21560-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21560-5.

Maniezzo:2008:LIO

[MBW08] Vittorio Maniezzo, Roberto Battiti, and Jean-Paul Watson, editors. Learning and Intelligent Optimization: Second International Conference, LION 2007 II, Trento, Italy, December 8–12, 2007. Selected Papers, volume 5313 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92694-1 (print), 3-540-92695-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92695-5.

Monien:1989:SAS

[MC89] B. Monien and Robert Cori, editors. STACS 89: 6th Annual Symposium on Theoretical Aspects of Computer Science, Paderborn, FRG, February 16–18, 1989, proceedings, volume 349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-50840-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.349.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0349.
htm;   http://www.springerlink.
com/content/978-0-387-
50840-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=349.

“Organized jointly in cooperation with EATCS by the Gesellschaft für Informatik (GI) and the Association française pour la cybernetique économique et technique (AFCET)”–Foreword.

Melham:1994:HOL

[MC94] T. F. Melham and Juanito Camilleri, editors. Higher order logic theorem proving and its applications: 7th International Workshop, Valletta, Malta, September 19–22, 1994: proceedings, volume 859 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58450-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 H54 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0859.
htm;   http://www.springerlink.
com/content/978-3-540-
58450-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=859.

Mugnier:1998:CST

[MC98] Marie-Laure Mugnier and Michel Chein, editors. Conceptual Structures: Theory, Tools and Applications: 6th International Conference on Conceptual Structures, ICCS’98, Montpellier, France, August 1998. Proceedings, volume 1453 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64791-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA387.2 .I58 1998. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1453.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64791-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1453.

Myers:2005:ABI

[MC05] Gene Myers and Rita Casadio, editors. Algorithms in bioinformatics: 5th international workshop, WABI 2005, Mallorca, Spain, October 3–6, 2005: proceedings, volume 3692 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29008-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R858.A2 W33 2005; QH324.2 .W33 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29008-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3692.

McAllester:2000:ADC

[McA00] David A. McAllester, editor. Automated deduction — CADE-17: 17th International Conference on Automated Deduction, Pittsburgh, PA, USA, June 17–20, 2000:

Proceedings, volume 1831 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67664-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1831. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1831.htm;
 http://www.springerlink.
com/content/978-3-540-
67664-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1831.

Melin:2007:FFL

[MCA^{+}07] Patricia Melin, Oscar Castillo, Luis T. Aguilar, Janusz Kacprzyk, and Witold Pedrycz, editors. Foundations of Fuzzy Logic and Soft Computing: 12th International Fuzzy Systems Association World Congress, IFSA 2007, Cancun, Mexico, June 18–21, 2007. Proceedings, volume 4529 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72917-8 (print), 3-540-72950-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72950-1.

Ma:2008:CNG

[MCA08] Yan Ma, Deokjai Choi, and Shingo Ata, editors. Challenges for Next Generation Network Operations and Service Management: 11th Asia-Pacific Network Operations and Management Symposium, APNOMS 2008, Beijing, China, October 22–24, 2008. Proceedings, volume 5297 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88622-2 (print), 3-540-88623-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88623-5.

McCalla:1996:AAI

[McC96] Gordon McCalla, editor. Advances in artificial intelligence: 11th Biennial Conference of the Canadian Society for Computational Studies of Intelligence, AI’96 Toronto, Ontario, Canada, May 21–24, 1996: proceedings, volume 1081 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61291-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .C36 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1081.htm;
http://www.springerlink.
com/content/978-3-540-
61291-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1081.

The AI 96 Conference is the 11th in a series of biennial artificial intelligence conferences sponsored by the Canadian Society for Computational Studies of Intelligence. …

McCune:1997:ADC

[McC97] W. (William) McCune, editor. Automated deduction—CADE-14: 14th International Conference on Automated Deduction, Townsville, North Queensland, Australia, July 13–17, 1997: proceedings, volume 1249 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63104-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96I57 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1249.
htm;   http://www.springerlink.
com/content/978-3-540-
63104-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1249.

Mueller:2008:OSM

[MCdS^{+}08] Matthias S. Mueller, Barbara M. Chapman, Bronis R. de Supinski, Allen D. Malony, and Michael Voss, editors. OpenMP Shared Memory Parallel Programming: International Workshops, IWOMP 2005 and IWOMP 2006, Eugene, OR, USA, June 1–4, 2005, Reims, France, June 12–15, 2006. Proceedings, volume 4315 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68554-5 (print), 3-540-68555-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68555-5.

Muller-Clostermann:2010:MME

[MCER10] Bruno Müller-Clostermann, Klaus Echtle, and Erwin P. Rathgeb, editors. Measurement, Modelling, and Evaluation of Computing Systems and Dependability and Fault Tolerance: 15th International GI/ITG Conference, MMB&DFT 2010, Essen, Germany, March 15–17, 2010. Proceedings, volume 5987 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12103-9 (print), 3-642-12104-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12104-3.

McGeoch:2008:EAI

[McG08] Catherine C. McGeoch, editor. Experimental Algorithms: 7th International Workshop, WEA 2008 Provincetown, MA, USA, May 30–June 1, 2008 Proceedings, volume 5038 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68548-0 (print), 3-540-68552-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68552-4.

McKay:1999:SEL

[McK99] Bob McKay, editor. Simulated evolution and learning: Second Asia-Pacific Conference on Simulated Evolution and Learning, SEAL ’98, Canberra, Australia, November 24–27, 1998: Selected Papers, volume 1585 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65907-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65 S42 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1585.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65907-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1585.

Madria:2006:DCI

[MCK^{+}06] Sanjay K. Madria, Kajal T. Claypool, Rajgopal Kannan, Prem Uppuluri, and Manoj Madhava Gore, editors. Distributed Computing and Internet Technology: Third International Conference, ICDCIT 2006, Bhubaneswar, India, December 20–23, 2006. Proceedings, volume 4317 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68379-8 (print), 3-540-68380-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68380-3.

Maglaveras:2006:BMD

[MCKB06] Nicos Maglaveras, Ioanna Chouvarda, Vassilis Koutkias, and Rüdiger Brause, editors. Biological and Medical Data Analysis: 7th International Symposium, ISBMDA 2006, Thessaloniki, Greece, December 7–8, 2006. Proceedings, volume 4345 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68063-2 (print), 3-540-68065-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68065-9.

Mira:1993:NTN

[MCP93] J. Mira, J. Cabestany, and A. Prieto, editors. New trends in neural computation: International Workshop on Artificial Neural Networks, IWANN ’93, Sitges, Spain, June 9–11, 1993: proceedings, volume 686 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56798-4 (Berlin), 0-387-56798-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I586 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0686.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        56798-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=686.

Mewhort:2010:HPC

[MCSN10] Douglas J. K. Mewhort, Natalie M. Cann, Gary W. Slater, and Thomas J. Naughton, editors. High Performance Computing Systems and Applications: 23rd International Symposium, HPCS 2009, Kingston, ON, Canada, June 14–17, 2009, Revised Selected Papers, volume 5976 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12658-8 (print), 3-642-12659-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12659-8.

Minden:2007:ANI

[MCSY07] Gary J. Minden, Kenneth L. Calvert, Marcin Solarski, and Miki Yamamoto, editors. Active Networks: IFIP TC6 6th International Working Conference, IWAN 2004, Lawrence, KS, USA, October 27–29, 2004. Revised Papers, volume 3912 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71499-5 (print), 3-540-71500-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71500-9.

Miclet:1996:GIL

[MD96] Laurent Miclet and Colin De la Higuera, editors. Grammatical inference: learning syntax from sentences: Third International Colloquium, ICGI-96, Montpellier, France, September 25–27, 1996: proceedings, volume 1147 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61778-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38G7 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1147.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61778-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1147.

Moreira:1999:OOT

[MD99] Ana M. D. Moreira and Serge Demeyer, editors. Object-oriented technology: ECOOP’99 workshop reader: ECOOP’99 workshops, panels, and posters, Lisbon, Portugal, June 14–18, 1999: Proceedings, volume 1743 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66954-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1999b; QA76.9.O35 E33 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1743.htm;
 http://www.springerlink.
com/content/978-3-540-
66954-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1743.

Murugesan:2001:WEM

[MD01] San Murugesan and Yogesh Deshpande, editors. Web engineering: managing diversity and complexity of Web application development, volume 2016 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42130-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .W3725 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2016.

                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        42130-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2016.

Marinai:2004:DAS

[MD04] Simone Marinai and Andreas Dengel, editors. Document analysis systems VI: 6th international workshop, DAS 2004, Florence, Italy, September 8–10, 2004: proceedings, volume 3163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23060-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1650 .D37 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        23060-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3163;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100557.

Moreno-Diaz:2001:CAS

[MDBF01] Roberto Moreno-Diaz, Bruno Buchberger, and Jose-Luis Freire, editors. Computer aided systems theory—EUROCAST 2001: a selection of papers from the 8th International Workshop on Computer Aided Systems Theory, Las Palmas de Gran Canaria, Spain, February 19–23, 2001: Revised Papers, volume 2178 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42959-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2178; TA345 .I62 2001. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2178.
htm;   http://www.springerlink.
com/content/978-3-540-
42959-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2178.

Meersman:2009:MMIa

[MDH09a] Robert Meersman, Tharam Dillon, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems: OTM 2009: Confederated International Conferences, CoopIS, DOA, IS, and ODBASE 2009, Vilamoura, Portugal, November 1–6, 2009, Proceedings, Part I, volume 5870 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05147-2 (print), 3-642-05148-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05148-7.

Meersman:2009:MMIb

[MDH09b] Robert Meersman, Tharam Dillon, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems: OTM 2009: Confederated International Conferences, CoopIS, DOA, IS, and ODBASE 2009, Vilamoura, Portugal, November 1–6, 2009, Proceedings, Part II, volume 5871 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05150-2 (print), 3-642-05151-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05151-7.

Meersman:2010:MMIa

[MDH10a] Robert Meersman, Tharam Dillon, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems: OTM 2010: Confederated International Conferences: CoopIS, IS, DOA and ODBASE, Hersonissos, Crete, Greece, October 25–29, 2010, Proceedings, Part I, volume 6426 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16933-3 (print), 3-642-16934-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16934-2.

Meersman:2010:MMIb

[MDH10b] Robert Meersman, Tharam Dillon, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems, OTM 2010: Confederated International Conferences: CoopIS, IS, DOA and ODBASE, Hersonissos, Crete, Greece, October 25–29, 2010, Proceedings, Part II, volume 6427 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16948-1 (print), 3-642-16949-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16949-6.

Meersman:2010:MMIc

[MDH10c] Robert Meersman, Tharam Dillon, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems: OTM 2010 Workshops: Confederated International Workshops and Posters: International Workshops: AVYTAT, ADI, DATAVIEW, EI2N, ISDE, MONET, OnTo-Content, ORM, P2P-CDVE, SeDeS, SWWS and OTMA. Hersonissos, Crete, Greece, October 25–29, 2010. Proceedings, volume 6428 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16960-0 (print), 3-642-16961-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16961-8.

Meersman:2011:MMI

[MDH11a] Robert Meersman, Tharam Dillon, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems:

OTM 2011 Workshops: Confederated International Workshops and Posters: EI2N+NSF ICE, ICSP+INBAST, ISDE, ORM, OTMA, SWWS+MONET+SeDeS, and VADER 2011, Hersonissos, Crete, Greece, October 17–21, 2011. Proceedings, volume 7046 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25125-0 (print), 3-642-25126-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .O86 2011. URL http:

//www.springerlink.
com/content/978-3-642-
25126-9.

Meersman:2011:MMIa

[MDH^{+}11b] Robert Meersman, Tharam Dillon, Pilar Herrero, Akhil Kumar, Manfred Reichert, Li Qing, Beng-Chin Ooi, Ernesto Damiani, Douglas C. Schmidt, Jules White, Manfred Hauswirth, Pascal Hitzler, and Mukesh Mohania, editors. On the Move to Meaningful Internet Systems: OTM 2011: Confederated International Conferences: CoopIS, DOA-SVI, and ODBASE 2011, Hersonissos, Crete, Greece, October 17–21, 2011, Proceedings, Part I, volume 7044 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25108-0 (print), 3-642-25109-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25109-2.

Meersman:2011:MMIb

[MDH^{+}11c] Robert Meersman, Tharam Dillon, Pilar Herrero, Akhil Kumar, Manfred Reichert, Li Qing, Beng-Chin Ooi, Ernesto Damiani, Douglas C. Schmidt, Jules White, Manfred Hauswirth, Pascal Hitzler, and Mukesh Mohania, editors. On the Move to Meaningful Internet Systems: OTM 2011: Confederated International Conferences: CoopIS, DOA-SVI, and ODBASE 2011, Hersonissos, Crete, Greece, October 17–21, 2011, Proceedings, Part II, volume 7045 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25105-6 (print), 3-642-25106-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25106-1.

Madabhushi:2011:PCI

[MDHB11] Anant Madabhushi, Jason Dowling, Henkjan Huisman, and Dean Barratt, editors. Prostate Cancer Imaging. Image Analysis and Image-Guided Interventions: International Workshop, Held in Conjunction with MICCAI 2011, Toronto, Canada, September 22, 2011. Proceedings, volume 6963 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23943-9 (print), 3-642-23944-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23944-1.

Marques:2005:PRIa

[MdlBP05a] Jorge S. Marques, Nicolás Pérez de la Blanca, and Pedro Pina, editors. Pattern Recognition and Image Analysis: Second Iberian Conference, IBPRIA 2005, Estoril, Portugal, June 7– 9, 2005, proceedings, Part I, volume 3522 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26153-2 (print), 3-540-32237-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1650 .I27 2005. URL http:

//www.springerlink.
com/content/978-3-540-
32237-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3522;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136825.

Marques:2005:PRIb

[MdlBP05b] Jorge S. Marques, Nicolás Pérez de la Blanca, and Pedro Pina, editors. Pattern Recognition and Image Analysis: Second Iberian Conference, IBPRIA 2005, Estoril, Portugal, June 7– 9, 2005, proceedings, Part II, volume 3523 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26154-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26154-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3523;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136831.

Moreno-Daaz:2004:CAS

[MDP04] Roberto Moreno-Dàaz and Franz Pichler, editors. Computer Aided Systems Theory—EUROCAST 2003 9th International Workshop on Computer Aided Systems Theory, Las Palmas de Gran Canaria, Spain, February 24–28, 2003: revised Selected Papers, volume 2809 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20221-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .I583 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2809.htm;
http://www.springerlink.
com/content/978-3-540-
20221-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2809;
http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b13239.

Mira:1998:IEA

[MdPA98] Jose Mira, Angel Pasqual del Pobil, and Moonis Ali, editors. Industrial and engineering applications of artificial intelligence and expert systems: 11th International Conference IEA/AIE-98, Benicassim, Castellon, Spain, June 1998: Proceedings. Part I. Methodology and Tools in Knowledge-Based Systems, volume 1415 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64582-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1415-1416. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1415.htm;
 http://www.springerlink.
com/content/978-3-540-
64582-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1415.

Contents: v. 1. Methodology and tools in knowledge-based systems / Jose Mira, Angel Pasqual del Pobil, Moonis Ali, eds. — v. 2. Tasks and methods in applied artificial intelligence / Angel Pasqual del Pobil, Jose Mira, Moonis Ali, eds.

Moreno-Diaz:2009:CAS

[MDPQA09] Roberto Moreno-Díaz, Franz Pichler, and Alexis Quesada-Arencibia, editors. Computer Aided Systems Theory — EU-ROCAST 2009: 12th International Conference, Las Palmas de Gran Canaria, Spain, February 15–20, 2009, Revised Selected Papers, volume 5717 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04771-8 (print), 3-642-04772-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04772-5.

Moreno-Diaz:2012:CASa

[MDPQA12a] Roberto Moreno-Díaz, Franz Pichler, and Alexis Quesada-Arencibia, editors. Computer Aided Systems Theory — EU-ROCAST 2011: 13th International Conference, Las Palmas de Gran Canaria, Spain, February 6–11, 2011, Revised Selected Papers, Part I, volume 6927 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27548-6 (print), 3-642-27549-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27549-4.

Moreno-Diaz:2012:CASb

[MDPQA12b] Roberto Moreno-Díaz, Franz Pichler, and Alexis Quesada-Arencibia, editors. Computer Aided Systems Theory — EU-ROCAST 2011: 13th International Conference, Las Palmas de Gran Canaria, Spain, February 6–11, 2011, Revised Selected Papers, Part II, volume 6928 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27578-8 (print), 3-642-27579-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27579-1.

Markopoulos:2008:FGS

[MdRIR08] Panos Markopoulos, Boris de Ruyter, Wijnand IJsselsteijn, and Duncan Rowland, editors. Fun and Games: Second International Conference, Eindhoven, The Netherlands, October 20– 21, 2008. Proceedings, volume 5294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88321-5 (print), 3-540-88322-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88322-7.

Muller:2009:EOA

[MdSC09] Matthias S. Müller, Bronis R. de Supinski, and Barbara M. Chapman, editors. Evolving OpenMP in an Age of Extreme Parallelism: 5th International Workshop on OpenMP, IWOMP 2009 Dresden, Germany, June 3–5, 2009 Proceedings, volume 5568 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02284-7 (print), 3-642-02303-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02303-3.

Mason:2007:MTP

[MDY07] Lorne Mason, Tadeusz Drwiega, and James Yan, editors. Managing Traffic Performance in Converged Networks: 20th International Teletraffic Congress, ITC20 2007, Ottawa, Canada, June 17–21, 2007. Proceedings, volume 4516 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72989-5 (print), 3-540-72990-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72990-7.

Madabhushi:2010:PCI

[MDY^{+}10] Anant Madabhushi, Jason Dowling, Pingkun Yan, Aaron Fenster, Purang Abolmaesumi, and Nobuhiko Hata, editors. Prostate Cancer Imaging. Computer-Aided Diagnosis, Prognosis, and Intervention: International Workshop, Held in Conjunction with MICCAI 2010, Beijing,China, September 24, 2010. Proceedings, volume 6367 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15988-5 (print), 3-642-15989-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15989-3.

Markopoulos:2004:AIS

[MEAC04] Panos Markopoulos, Berry Eggen, Emile Aarts, and James L. Crowley, editors. Ambient Intelligence: Second European Symposium, EUSAI 2004, Eindhoven, The Netherlands, November 8–11, 2004. Proceedings, volume 3295 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23721-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3295.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        23721-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3295;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b102265.

Medeiros:2005:AST

[MEB05] Claudia Bauzer Medeiros, Max J. Egenhofer, and Elisa Bertino, editors. Advances in Spatial and Temporal Databases: 9th International Symposium, SSTD 2005, Angra dos Reis, Brazil, August 22–24, 2005: Proceedings, volume 3633 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28127-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S6925 2005. URL http://

                                        springerlink.metapress.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3633;   http:
                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28127-6.

Mehlhorn:1985:SAS

[Meh85] Kurt Mehlhorn, editor. STACS 85: 2nd Annual Symposium on Theoretical Aspects of Computer Science, Saarbrucken, January 3–5, 1985, volume 182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-13912-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.182. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0182.htm;
 http://www.springerlink.
com/content/978-0-387-
13912-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=182.

Meinel:1989:MBP

[Mei89] Christoph Meinel. Modified branching programs and their computational power, volume 370 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51340-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 132 pp. LCCN QA274.76 .M451 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0370.htm;
 http://www.springerlink.
com/content/978-0-387-
51340-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=370.

Mei:2008:HCS

[Mei08] Hong Mei, editor. High Confidence Software Reuse in Large Systems: 10th International Conference on Software Reuse, ICSR 2008, Beijing, China, May 25–29, 2008 Proceedings, volume 5030 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68062-4 (print), 3-540-68073-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68073-4.

Manolopoulos:2003:AIP

[MEK03] Yannis Manolopoulos, Skevos Evripidou, and Antonis C. Kakas, editors. Advances in informatics: 8th Panhellenic Conference on Informatics, PCI 2001, Nicosia, Cyprus, November 8–10, 2001: revised Selected Papers, volume 2563 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-07544-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T58.5 .P35 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2563.
htm;   http://www.springerlink.
com/content/978-3-540-
07544-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2563.

Melton:1986:MFP

[Mel86] Austin Melton, editor.

Mathematical foundations of programming semantics: international conference, Manhattan, Kansas, April 11–12, 1985: proceedings, volume 239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16816-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .M381 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0239.htm;
http://www.springerlink.
com/content/978-0-387-
16816-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=239.

Melnik:2004:GMM

[Mel04] Sergey Melnik. Generic Model Management: Concepts and Algorithms, volume 2967 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21980-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xvii + 238 pp. LCCN QA76.6245 .M45 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2967.htm;
 http://www.springerlink.
com/content/978-3-540-
21980-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2967;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97859.

Mellouli:2007:SQA

[Mel07] Khaled Mellouli, editor. Symbolic and Quantitative Approaches to Reasoning with Uncertainty: 9th European Conference, ECSQARU 2007, Hammamet, Tunisia, October 31 — November 2, 2007. Proceedings, volume 4724 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75255-2 (print), 3-540-75256-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75256-1.

Menezes:2005:TCC

[Men05] Alfred Menezes, editor.

Topics in cryptology CT-RSA 2005: The cryptographers’ track at the RSA Conference 2005, San Francisco, CA, USA, February 14–18, 2005. Proceedings, volume 3376 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24399-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 R753 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24399-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3376;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105222.

Menezes:2007:ACC

[Men07] A. J. (Alfred J.) Menezes, editor. Advances in cryptology — CRYPTO 2007: 27th Annual International Cryptology Conference, Santa Barbara, CA, USA, August 19–23, 2007: proceedings, volume 4622 of Lecture notes in computer science, 0302-9743. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-74142-9 (paperback). LCCN QA76.9.A25 C79 2007. URL http:

//www.springerlink.
com/content/978-3-540-
74142-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=4622.

Meyer:1995:CPS

[Mey95] Manfred Meyer. Constraint processing: Selected Papers, volume 923 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 0-387-59479-5, 3-540-59479-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). 287 pp. LCCN QA76.612 .C68 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0923.htm;
 http://www.springerlink.
com/content/978-0-387-
59479-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=923.

Papers from two workshops held in 1993 and 1994.

MeyeraufderHeide:2001:AEA

[Mey01] F. Meyer auf der Heide, editor. Algorithms-ESA 2001: 9th annual European symposium, Aarhus, Denmark, August 28– 31, 2001: Proceedings, volume 2161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42493-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 2001; QA267.A1 L43 no.2161. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2161.htm;
 http://www.springerlink.
com/content/978-3-540-
42493-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2161.

Mezini:2011:EOO

[Mez11] Mira Mezini, editor.

ECOOP 2011 — Object-Oriented Programming: 25th European Conference, Lancaster, Uk, July 25–29, 2011 Proceedings, volume 6813 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22654-X (print), 3-642-22655-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22655-7.

Marsland:2001:CGS

[MF01] T. Marsland and I. Frank, editors. Computers and Games: Second International Conference, CG 2000, Hamamatsu, Japan, October 26–28, 2000: Revised Papers, volume 2063 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43080-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .C4166 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2063.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43080-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2063.

Monien:2002:EPP

[MF02] Burkhard Monien and Rainer Feldmann, editors. Euro-Par 2002, parallel processing: 8th International Euro-Par Conference, Paderborn, Germany, August 27– 30, 2002: Proceedings, volume 2400 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44049-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I553 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2400.htm;
 http://www.springerlink.
com/content/978-3-540-
44049-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2400.

Moreau:2006:PAD

[MF06] Luc Moreau and Ian Foster, editors. Provenance and Annotation of Data: International Provenance and Annotation Workshop, IPAW 2006, Chicago, IL, USA, May 3–5, 2006, Revised Selected Papers, volume 4145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46302-X (print), 3-540-46303-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46303-0.

Matsuura:2008:AIC

[MF08] Kanta Matsuura and Eiichiro Fujisaki, editors. Advances in Information and Computer Security: Third International Workshop on Security, IWSEC 2008, Kagawa, Japan, November 25–27, 2008. Proceedings, volume 5312 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89597-3 (print), 3-540-89598-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89598-5.

Maino:2011:IAPa

[MF11a] Giuseppe Maino and Gian Luca Foresti, editors. Image Analysis and Processing — ICIAP 2011: 16th International Conference, Ravenna, Italy, September 14–16, 2011, Proceedings, Part I, volume 6978 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24084-4 (print), 3-642-24085-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24085-0.

Maino:2011:IAPb

[MF11b] Giuseppe Maino and Gian Luca Foresti, editors. Image Analysis and Processing — ICIAP 2011: 16th International Conference, Ravenna, Italy, September 14–16, 2011, Proceedings, Part II, volume 6979 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24087-9 (print), 3-642-24088-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24088-1.

Mester:2011:PRD

[MF11c] Rudolf Mester and Michael Felsberg, editors. Pattern Recognition: 33rd DAGM Symposium, Frankfurt/Main, Germany, August 31 — September 2, 2011. Proceedings, volume 6835 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23122-5 (print), 3-642-23123-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23123-0.

Mira:2009:BAA

[MFÁ^{+}09a] José Mira, José Manuel Ferrández, José R. Álvarez, Félix de la Paz, and F. Javier Toledo, editors. Bioinspired Applications in Artificial and Natural Computation: Third International Work-Conference on the Interplay Between Natural and Artificial Computation, IWINAC 2009, Santiago de Compostela, Spain, June 22– 26, 2009, Proceedings, Part II, volume 5602 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02266-9 (print), 3-642-02267-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02267-8.

Mira:2009:MMA

[MFÁ^{+}09b] José Mira, José Manuel Ferrández, José R. Álvarez, Félix de la Paz, and F. Javier Toledo, editors. Methods and Models in Artificial and Natural Computation. A Homage to Professor Mira’s Scientific Legacy: Third International Work-Conference on the Interplay Between Natural and Artificial Computation, IWINAC 2009, Santiago de Compostela, Spain, June 22– 26, 2009, Proceedings, Part I, volume 5601 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02263-4 (print), 3-642-02264-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02264-7.

Meadows:2012:STM

[MFG12] Catherine Meadows and Carmen Fernandez-Gago, editors. Security and Trust Management: 7th International Workshop, STM 2011, Copenhagen, Denmark, June 27–28, 2011, Revised Selected Papers, volume 7170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29962-8 (print), 3-642-29963-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29963-6.

McDaniel:2007:ISS

[MG07a] Patrick McDaniel and Shyam K. Gupta, editors. Information Systems Security: Third International Conference, ICISS 2007, Delhi, India, December 16–20, 2007. Proceedings, volume 4812 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77085-2 (print), 3-540-77086-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77086-2.

Moreira:2007:EAC

[MG07b] Ana Moreira and John Grundy, editors. Early Aspects: Current Challenges and Future Directions: 10th International Workshop, Vancouver, Canada, March 13, 2007, Revised Selected Papers, volume 4765 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76810-6 (print), 3-540-76811-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76811-1.

Maicher:2008:STM

[MG08] Lutz Maicher and Lars Marius Garshol, editors. Scaling Topic Maps: Third International Conference on Topic Maps Research and Applications, TMRA 2007 Leipzig, Germany, October 11–12, 2007 Revised Selected Papers, volume 4999 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70873-1 (print), 3-540-70874-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70874-2.

Mirandola:2009:AAS

[MGH09] Raffaela Mirandola, Ian Gorton, and Christine Hofmeister, editors. Architectures for Adaptive Software Systems: 5th International Conference on the Quality of Software Architectures, QoSA 2009, East Stroudsburg, PA, USA, June 24–26, 2009 Proceedings, volume 5581 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02350-9 (print), 3-642-02351-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02351-4.

Murgante:2011:CSAa

[MGI^{+}11a] Beniamino Murgante, Osvaldo Gervasi, Andrés Iglesias, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2011: International Conference, Santander, Spain, June 20–23, 2011. Proceedings, Part I, volume 6782 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21927-6 (print), 3-642-21928-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21928-3.

Murgante:2011:CSAb

[MGI^{+}11b] Beniamino Murgante, Osvaldo Gervasi, Andrés Iglesias, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2011: International Conference, Santander, Spain, June 20–23, 2011. Proceedings, Part II, volume 6783 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21886-5 (print), 3-642-21887-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21887-3.

Murgante:2011:CSAc

[MGI^{+}11c] Beniamino Murgante, Osvaldo Gervasi, Andrés Iglesias, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2011: International Conference, Santander, Spain, June 20–23, 2011. Proceedings, Part III, volume 6784 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21930-6 (print), 3-642-21931-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21931-3.

Murgante:2011:CSAd

[MGI^{+}11d] Beniamino Murgante, Osvaldo Gervasi, Andrés Iglesias, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2011: International Conference, Santander, Spain, June 20–23, 2011. Proceedings, Part IV, volume 6785 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21897-0 (print), 3-642-21898-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21898-9.

Murgante:2011:CSAe

[MGI^{+}11e] Beniamino Murgante, Osvaldo Gervasi, Andrés Iglesias, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2011: International Conference, Santander, Spain, June 20–23, 2011. Proceedings, Part V, volume 6786 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21933-0 (print), 3-642-21934-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21934-4.

Murgante:2012:CSAa

[MGM^{+}12a] Beniamino Murgante, Osvaldo Gervasi, Sanjay Misra, Nadia Nedjah, Ana Maria A. C. Rocha, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2012: 12th International Conference, Salvador de Bahia, Brazil, June 18– 21, 2012, Proceedings, Part I, volume 7333 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31124-5 (print), 3-642-31125-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31125-3.

Murgante:2012:CSAb

[MGM^{+}12b] Beniamino Murgante, Osvaldo Gervasi, Sanjay Misra, Nadia Nedjah, Ana Maria A. C. Rocha, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2012: 12th International Conference, Salvador de Bahia, Brazil, June 18– 21, 2012, Proceedings, Part II, volume 7334 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31074-5 (print), 3-642-31075-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31075-1.

Murgante:2012:CSAc

[MGM^{+}12c] Beniamino Murgante, Osvaldo Gervasi, Sanjay Misra, Nadia Nedjah, Ana Maria A. C. Rocha, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2012: 12th International Conference, Salvador de Bahia, Brazil, June 18– 21, 2012, Proceedings, Part III, volume 7335 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31136-9 (print), 3-642-31137-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31137-6.

Murgante:2012:CSAd

[MGM^{+}12d] Beniamino Murgante, Osvaldo Gervasi, Sanjay Misra, Nadia Nedjah, Ana Maria A. C. Rocha, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2012: 12th International Conference, Salvador de Bahia, Brazil, June 18– 21, 2012, Proceedings, Part IV, volume 7336 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31127-X (print), 3-642-31128-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31128-4.

Muller:2012:MCB

[MGSM12] Henning Müller, Hayit Greenspan, and Tanveer Syeda-Mahmood, editors. Medical Content-Based Retrieval for Clinical Decision Support: Second MICCAI International Workshop, MCBR-CDS 2011, Toronto, ON, Canada, September 22, 2011, Revised Selected Papers, volume 7075 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28459-0 (print), 3-642-28460-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28460-1.

Melnik:2004:ALS

[MH04] Grigori Melnik and Harald Holz, editors. Advances in Learning Software Organizations: 6th International Workshop, LSO 2004, Banff, Canada, June 20–21, 2004: Proceedings, volume 3096 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22192-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758.L76 2004. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3096.
htm;   http://www.springerlink.
com/content/978-3-540-
22192-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3096;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98340.

McLysaght:2005:CGR

[MH05] Aoife McLysaght and Daniel H. Huson, editors. Comparative genomics: RECOMB 2005 international workshop, RCG 2005, Dublin, Ireland, September 18–20, 2005: proceedings, volume 3678 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28932-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH447 .R34 2005. URL http:

//www.springerlink.
com/content/978-3-540-
28932-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3678.

Merz:2011:ECC

[MH11] Peter Merz and Jin-Kao Hao, editors. Evolutionary Computation in Combinatorial Optimization: 11th European Conference, Evo-COP 2011, Torino, Italy, April 27–29, 2011. Proceedings, volume 6622 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20363-9 (print), 3-642-20364-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20364-0.

Mertsching:2009:KAA

[MHA09] Bärbel Mertsching, Marcus Hund, and Zaheer Aziz, editors. KI 2009: Advances in Artificial Intelligence: 32nd Annual German Conference on AI, Paderborn, Germany, September 15– 18, 2009. Proceedings, volume 5803 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04616-9 (print), 3-642-04617-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04617-9.

Meersman:2009:MMIc

[MHD09] Robert Meersman, Pilar Herrero, and Tharam Dillon, editors. On the Move to Meaningful Internet Systems: OTM 2009 Workshops: Confederated International Workshops and Posters, ADI, CAMS, EI2N, ISDE, IWSSA, MONET, OnToContent, ODIS, ORM, OTM Academy, SWWS, SEMELS, Beyond SAWSDL, and COMBEK 2009, Vilamoura, Portugal, November 1–6, 2009. Proceedings, volume 5872 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05289-4 (print), 3-642-05290-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05290-3.

Miksch:2005:AIM

[MHK05] Silvia Miksch, J. (Jim) Hunter, and E. T. Keravnou, editors. Artificial intelligence in medicine: 10th Conference on Artificial Intelligence in Medicine, AIME 2005, Aberdeen, UK, July 23–27, 2005: proceedings, volume 3581 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27831-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R859.7.A78; R859.7.A78 C66 2005eb; R859.7.A78 C66 2005; Internet. URL

                                        http://www.springerlink.
                                        com/content/978-3-540-
                                        27831-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3581.

Muller-Hannemann:2010:AEB

[MHS10] Matthias Müller-Hannemann and Stefan Schirra, editors. Algorithm Engineering: Bridging the Gap between Algorithm Theory and Practice, volume 5971 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14865-4 (print), 3-642-14866-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14866-8.

Morzy:2012:ADI

[MHW12] Tadeusz Morzy, Theo Härder, and Robert Wrembel, editors. Advances in Databases and Information Systems: 16th East European Conference, ADBIS 2012, Poznań, Poland, September 18–21, 2012. Proceedings, volume 7503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33073-8 (print), 3-642-33074-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33074-2.

Michalewicz:1990:SSD

[Mic90] Zbigniew Michalewicz, editor. Statistical and scientific database management: fifth international conference, V SS-DBM, Charlotte, N.C., USA, April 3–5, 1990: proceedings, volume 420 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-52342-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I583 1990. DM42.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0420.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        52342-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=420.

Sponsored by the University of North Carolina at Charlotte.

Micciancio:2010:TCT

[Mic10] Daniele Micciancio, editor. Theory of Cryptography: 7th Theory of Cryptography Conference, TCC 2010, Zurich, Switzerland, February 9– 11, 2010. Proceedings, volume 5978 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11798-8 (print), 3-642-11799-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11799-2.

Middeldorp:2001:RTA

[Mid01] Aart Middeldorp, editor. Rewriting Techniques and Applications: 12th International Conference, RTA 2001, Utrecht, The Netherlands, May 22–24, 2001: Proceedings, volume 2051 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42117-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .R48 2001. URL http:

//link.springer.de/
link/service/series/
0558/tocs/t2051.htm;
http://www.springerlink.
com/content/978-3-540-
42117-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2051.

Milner:1980:CCS

[Mil80] R. (Robin) Milner. A calculus of communicating systems, volume 92 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10235-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 171 pp. LCCN QA267 .M53. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0092.htm;
 http://www.springerlink.
com/content/978-0-387-
10235-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=92.

Miller:2000:ESB

[Mil00] Julian Miller, editor. Evolvable systems: from biology to hardware: Third International Conference, ICES 2000, Edinburgh, Scotland, April 17–19, 2000: Proceedings, volume 1801 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67338-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1801; QA76.618 .I57 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1801.
htm;   http://www.springerlink.
com/content/978-3-540-

67338-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1801.

Miller:2001:GPE

[Mil01] Julian Miller, editor.

Genetic programming: 4th European Conference, EuroGP 2001, Lake Como, Italy, April 18– 20, 2001: Proceedings, volume 2038 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41899-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.623 .E9 2001; QA267.A1 L43 no.2038. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2038.htm;
http://www.springerlink.
com/content/978-3-540-
41899-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2038.

Milano:2012:PPC

[Mil12] Michela Milano, editor. Principles and practice of constraint programming: 18th International Conference, CP 2012, Québec City, QC, Canada, October 8– 12, 2012. Proceedings, volume 7514 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33557-8, 3-642-33558-6 (e-book). LCCN QA76.612 .C6 2012. URL

http://www.springerlink.
com/content/978-3-642-
33557-0;   http://www.
springerlink.com/content/
978-3-642-33558-7.

Miola:1990:DIS

[Mio90] A. Miola, editor. Design and implementation of symbolic computation systems: International Symposium DISCO ’90, Capri, Italy, April 10–12, 1990: proceedings, volume 429 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52531-9 (Berlin), 0-387-52531-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0429.htm;
 http://www.springerlink.
com/content/978-3-540-
52531-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=429.

Miola:1993:DIS

[Mio93] A. Miola, editor. Design and implementation of symbolic computation systems: International Symposium, DISCO ’93, Gmunden, Austria, September 15– 17, 1993: proceedings, volume 722 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57235-X (Berlin), 0-387-57235-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0722.htm;
http://www.springerlink.
com/content/978-3-540-
57235-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=722.

“Sponsored by the Research Institute for Symbolic Computation (University J. Kepler, Linz, Austria) and by the Dipartimento di informatica e sistemistica (University ‘La Sapienza’, Roma, Italy)”–Foreword.

Mittal:1998:ATA

[Mit98] Vibhu O. Mittal, editor. Assistive technology and artificial intelligence: applications in robotics, user interfaces, and natural language processing, volume 1458 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64790-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ211 .A85 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1458.
htm;   http://www.springerlink.
com/content/978-3-540-
64790-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1458.

Mittermeir:2005:CLI

[Mit05] Roland T. Mittermeir, editor. From Computer Literacy to Informatics Fundamentals: International Conference on Informatics in Secondary Schools — Evolution and Perspectives, ISSEP 2005: Klagenfurt, Austria, March 30– April 1, 2005: Proceedings, volume 3422 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25336-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.27 .I563 2005. URL http:

//www.springerlink.
com/content/978-3-540-
25336-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3422;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107112.

Mittermeir:2006:IEB

[Mit06] Roland T. Mittermeir, editor. Informatics Education — The Bridge between Using and Understanding Computers: International Conference in Informatics in Secondary Schools — Evolution and Perspectives, ISSEP 2006, Vilnius, Lithuania, November 7– 11, 2006. Proceedings, volume 4226 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48218-0 (print), 3-540-48227-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48227-7.

Maltoni:2004:BAE

[MJ04] Davide Maltoni and Anil K. Jain, editors. Biometric Authentication: ECCV 2004 International Workshop, BioAW 2004, Prague, Czech Republic, May 15th, 2004: Proceedings, volume 3087 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22499-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3E27 2004.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3087.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        22499-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3087;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99174.

Meyer:2007:SEA

[MJ07] Bertrand Meyer and Mathai Joseph, editors. Software Engineering Approaches for Offshore and Outsourced Development: First International Conference, SEAFOOD 2007, Zurich, Switzerland, February 5– 6, 2007. Revised Papers, volume 4716 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75541-1 (print), 3-540-75542-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75542-5.

Mutzel:2002:GDI

[MJL02] Petra Mutzel, M. Jünger, and Sebastian Leipert, editors. Graph drawing: 9th International Symposium, GB 2001, Vienna, Austria, September 23–26, 2001: Revised Papers, volume 2265 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43309-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2265.
htm;   http://www.springerlink.
com/content/978-3-540-
43309-5;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2265.

Masoodian:2004:CHI

[MJR04] Masood Masoodian, Steve Jones, and Bill Rogers, editors. Computer human interaction: 6th Asia Pacific conference, APCHI 2004, Rotorua, New Zealand, June 29– July 2, 2004: proceedings, volume 3101 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22312-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 A73 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        22312-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3101;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98382.

Ma:2006:UIC

[MJYT06] Jianhua Ma, Hai Jin, Laurence T. Yang, and Jeffrey J.-P. Tsai, editors. Ubiquitous Intelligence and Computing: Third International Conference, UIC 2006, Wuhan, China, September 3–6, 2006. Proceedings, volume 4159 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38091-4 (print), 3-540-38092-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38092-4.

Morrison:1996:ADB

[MK96] R. (Ronald) Morrison and Jessie Kennedy, editors. Advances in databases: 14th British National Conference on Databases, BNCOD 14, Edinburgh, Scotland, United Kingdom, July 3–5, 1996: proceedings, volume 1094 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61442-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B75 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1094.htm;
 http://www.springerlink.
com/content/978-3-540-
61442-5;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1094.

Mohnen:2001:IFL

[MK01] Markus Mohnen and Pieter Koopman, editors. Implementation of functional languages: 12th International Workshop, IFL 2000, Aachen, Germany, September 4– 7, 2000: Selected Papers, volume 2011 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41919-5 (paperback ). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2011; QA76.62 .I56 2000. URL http:

                                        //link.springer.de/
                                        link/service/series/
                                        0558/tocs/t2011.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41919-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2011.

Michaelis:2003:PRD

[MK03a] Bernd Michaelis and Gerald Krell, editors. Pattern Recognition: 25th DAGM Symposium, Magdeburg, Germany, September 10–12, 2003: Proceedings, volume 2781 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40861-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1650 .D35 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2781.htm;
 http://www.springerlink.
com/content/978-3-540-
40861-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2781;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12010.

Moro:2003:APP

[MK03b] Gianluca Moro and Manolis Koubarakis, editors. Agents and Peer-to-Peer Computing: First International Workshop, AP2PC 2002 Bologna, Italy, July 15, 2002 Revised and Invited Papers, volume 2530 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40538-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.525 .A615 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2530.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        40538-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2530.

Mossakowski:2012:RTA

[MK12] Till Mossakowski and Hans-Jörg Kreowski, editors. Recent Trends in Algebraic Development Techniques: 20th International Workshop, WADT 2010, Etelsen, Germany, July 1–4, 2010, Revised Selected Papers, volume 7137 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28411-6 (print), 3-642-28412-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28412-0.

Mokhtari:2009:AAH

[MKB^{+}09] Mounir Mokhtari, Ismail Khalil, Jérémy Bauchet, Daqing Zhang, and Chris Nugent, editors. Ambient Assistive Health and Wellness Management in the Heart of the City: 7th International Conference on Smart Homes and Health Telematics, ICOST 2009, Tours, France, July 1–3, 2009.

Proceedings, volume 5597 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02867-5 (print), 3-642-02868-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02868-7.

Magedanz:2005:MAT

[MKPV05] Thomas Magedanz, Ahmed Karmouch, Samuel Pierre, and Iakovos Venieris, editors. Mobility Aware Technologies and Applications: Second International Workshop, MATA 2005, Montréal, Canada, October 17–19, 2005. Proceedings, volume 3744 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29410-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29410-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3744.

Miesenberger:2012:CHPa

[MKPZ12a] Klaus Miesenberger, Arthur Karshmer, Petr Penaz, and Wolfgang Zagler, editors. Computers Helping People with Special Needs: 13th International Conference, IC-CHP 2012, Linz, Austria, July 11–13, 2012, Proceedings, Part I, volume 7382 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31521-6 (print), 3-642-31522-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31522-0.

Miesenberger:2012:CHPb

[MKPZ12b] Klaus Miesenberger, Arthur Karshmer, Petr Penaz, and Wolfgang Zagler, editors. Computers Helping People with Special Needs: 13th International Conference, IC-CHP 2012, Linz, Austria, July 11–13, 2012, Proceedings, Part II, volume 7383 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31533-X (print), 3-642-31534-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31534-3.

Mitrou:2004:NNT

[MKR^{+}04] Nikolas Mitrou, Kimon Kontovasilis, George N. Rouskas, Ilias Iliadis, and Lazaros Merakos, editors. NETWORKING 2004: Networking Technologies, Services, and Protocols: Performance of Computer and Communication Networks: Mobile and Wireless Communications: Third International IFIP-TC6 Networking Conference, Athens, Greece, May 9– 14, 2004: Proceedings, volume 3042 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21959-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .N46686 2004. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3042.
htm;   http://www.springerlink.
com/content/978-3-540-
21959-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3042;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b97826.

Miyaji:2007:AIC

[MKR07] Atsuko Miyaji, Hiroaki Kikuchi, and Kai Rannenberg, editors. Advances in Information and Computer Security: Second International Workshop on Security, IWSEC 2007, Nara, Japan, October 29–31, 2007. Proceedings, volume 4752 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CO-DEN LNCSD9. ISBN 3-540-75650-7 (print), 3-540-75651-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I573 2007. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75651-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=book\%26isbn=
                                        978-3-540-75650-7;
                                        http://www.springerlink.
                                        com/openurl.asp?genre=
                                        issue\%26issn=0302-
                                        9743\%26volume=4752.

Masunaga:1998:WCA

[MKT98] Y. (Yoshifumi) Masunaga, Takuya Katayama, and Michiharu Tsukamoto, editors. Worldwide computing and its applications — WWCA’98:

Second International Conference, Tsukuba, Japan, March 4–5, 1998: Proceedings, volume 1368 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64216-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5.W18 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1368.
htm;   http://www.springerlink.
com/content/978-3-540-
64216-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1368.

Miesenberger:2002:CHP

[MKZ02] Klaus Miesenberger, Joachim Klaus, and Wolfgang Zagler, editors. Computer Helping People with Special Needs: 8th International Conference, IC-CHP 2002, Linz, Austria, July 15–20, 2002. Proceedings, volume 2398 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43904-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HV1569.5 .I57 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2398.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43904-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2398.

Miesenberger:2004:CHP

[MKZB04] Klaus Miesenberger, Joachim Klaus, Wolfgang Zagler, and Dominique Burger, editors. Computer Helping People with Special Needs: 9th International Conference, IC-CHP 2004 Paris, France, July 7–9, 2004 Proceedings, volume 3118 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22334-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3118.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22334-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3118;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98426.

Miesenberger:2006:CHP

[MKZK06] Klaus Miesenberger, Joachim Klaus, Wolfgang L. Zagler, and Arthur I. Karshmer, editors. Computers Helping People with Special Needs: 10th International Conference, ICCHP 2006, Linz, Austria, July 11– 13, 2006. Proceedings, volume 4061 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36020-4 (print), 3-540-36021-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36021-6.

Miesenberger:2008:CHP

[MKZK08] Klaus Miesenberger, Joachim Klaus, Wolfgang Zagler, and Arthur Karshmer, editors. Computers Helping People with Special Needs: 11th International Conference, IC-CHP 2008, Linz, Austria, July 9–11, 2008. Proceedings, volume 5105 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70539-2 (print), 3-540-70540-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70540-6.

Miesenberger:2010:CHPa

[MKZK10a] Klaus Miesenberger, Joachim Klaus, Wolfgang Zagler, and Arthur Karshmer, editors. Computers Helping People with Special Needs: 12th International Conference, IC-CHP 2010, Vienna, Austria, July 14–16, 2010. Proceedings, volume 6179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14096-3 (print), 3-642-14097-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14097-6.

Miesenberger:2010:CHPb

[MKZK10b] Klaus Miesenberger, Joachim Klaus, Wolfgang Zagler, and Arthur Karshmer, editors. Computers Helping People with Special Needs: 12th International Conference, IC-CHP 2010, Vienna, Austria, July14-16, 2010, Proceedings, Part II, volume 6180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14099-8 (print), 3-642-14100-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14100-3.

Moore:1995:FPL

[ML95] Will R. Moore and Wayne Luk, editors. Field-programmable logic and applications: 5th International Workshop, FPL ’95, Oxford, United Kingdom, August 29– September 1, 1995: proceedings, volume 975 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60294-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.G36 I48 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0975.
htm;   http://www.springerlink.
com/content/978-3-540-
60294-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=975.

Mammeri:2004:HSN

[ML04] Zoubir Mammeri and Pascal Lorenz, editors. High Speed Networks and Multimedia Communications: 7th IEEE International Conference, HSNMC 2004, Toulouse, France, June 30–July 2, 2004: Proceedings, volume 3079 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22262-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3079.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22262-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3079;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98322.

Mauri:2011:DLT

[ML11] Giancarlo Mauri and Alberto Leporati, editors. Developments in Language Theory: 15th International Conference, DLT 2011, Milan, Italy, July 19–22, 2011. Proceedings, volume 6795 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22320-6 (print), 3-642-22321-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22321-1.

Meo:2004:DSD

[MLK04] Rosa Meo, Pier Luca Lanzi, and Mika Klemettinen, editors. Database Support for Data Mining Applications: Discovering Knowledge with Inductive Queries, volume 2682 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22479-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2682.htm;
 http://www.springerlink.
com/content/978-3-540-
22479-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2682;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b99016.

Martin-Lof:1990:CIC

[MLM90] Per Martin-Lof and G. E. Mints, editors. COLOG-88: International Conference on Computer Logic Tallinn, USSR, December 12–16, 1988: proceedings, volume 417 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52335-9 (Berlin), 0-387-52335-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 I57 1988. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0417.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        52335-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=417.

Menze:2011:MCV

[MLTC11] Bjoern Menze, Georg Langs, Zhuowen Tu, and Antonio Criminisi, editors. Medical Computer Vision. Recognition Techniques and Applications in Medical Imaging: International MIC-CAI Workshop, MCV 2010, Beijing, China, September 20, 2010, Revised Selected Papers, volume 6533 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18420-0 (print), 3-642-18421-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18421-5.

Marik:1993:DES

[MLW93] V. Marik, Jiri Lazansky, and Roland R. Wagner, editors. Database and expert systems applications, volume 720 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57234-1 (Berlin), 0-387-57234-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D32 D363 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0720.htm;
 http://www.springerlink.
com/content/978-3-540-
57234-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=720.

Proceedings of the Database and Expert Systems Applications Conference.

Messina:1982:PMM

[MM82] P. C. Messina and A. Murli, editors. Problems and methodologies in mathematical software production: international seminar held at Sorrento, Italy, November 3–8, 1980, volume 142 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11603-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .P76 1982. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0142.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        11603-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=142.

Papers presented at International Seminar on Problems and Methodologies in Mathematical Software Production held in Sorrento, Italy and sponsored by the National Committee for the Mathematical Sciences and the National Group for Mathematical Informatics.

Martins:1989:EPC

[MM89] Joao P. Martins and Ernesto J. M. Morgado, editors. EPIA 89: 4th Portuguese Conference on Artificial Intelligence, Lisbon, Portugal, September 26–29, 1989, proceedings, volume 390 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51665-4 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P67 1989. DM57.50.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0390.
htm;   http://www.springerlink.
com/content/978-0-387-
51665-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=390.

MeyeraufderHeide:1996:ALP

[MM96] F. (Friedhelm) Meyer auf der Heide and B. Monien, editors. Automata, languages and programming: 23rd international colloquium, ICALP ’96, Paderborn, Germany, July 8–12, 1996: proceedings, volume 1099 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61440-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I5783 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1099.htm;
 http://www.springerlink.
com/content/978-3-540-
61440-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1099.

Mora:1997:AAA

[MM97] Teo Mora and H. F. Mattson, editors. Applied algebra, algebraic algorithms, and errorcorrecting codes: 12th International Symposium, AAECC-12, Toulouse, France, June 23–27, 1997: proceedings, volume 1255 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63163-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1255.htm;
 http://www.springerlink.
com/content/978-3-540-

63163-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1255.

Margaria:2001:CHD

[MM01] Tiziana Margaria and T. F. Melham, editors. Correct hardware design and verification methods: 11th IFIP WG 10.5 Advanced Research Working Conference, CHARME 2001, Livingston, Scotland, UK, September 4– 7, 2001: Proceedings, volume 2144 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42541-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874 .A3353 2001; QA267.A1 L43 no.2144. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2144.htm;
http://www.springerlink.
com/content/978-3-540-
42541-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2144.

Matousek:2003:TSD

[MM03] Václav Matoušek and Pavel Mautner, editors. Text, Speech and Dialogue: 6th International Conference, TSD 2003, České Budějovice, Czech Republic, September 8–12, 2003: Proceedings, volume 2807 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20024-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 T48 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2807.
htm;   http://www.springerlink.
com/content/978-3-540-
20024-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2807;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13236.

Meziane:2004:NLP

[MM04] Farid Meziane and Elisabeth Métais, editors. Natural Language Processing and Information Systems: 9th International Conference on Applications of Natural Language to Information Systems, NLDB 2004, Salford, UK, June 23–25, 2004: Proceedings, volume 3136 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22564-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3136.htm;
http://www.springerlink.
com/content/978-3-540-
22564-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3136;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98754.

Matousek:2007:TSD

[MM07] Václav Matoušek and Pavel Mautner, editors. Text, Speech and Dialogue: 10th International Conference, TSD 2007, Pilsen, Czech Republic, September 3–7, 2007. Proceedings, volume 4629 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74627-7 (print), 3-540-74628-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74628-7.

Marchiori:2008:ECM

[MM08] Elena Marchiori and Jason H. Moore, editors. Evolutionary Computation, Machine Learning and Data Mining in Bioinformatics: 6th European Conference, Evo-BIO 2008, Naples, Italy, March 26–28, 2008. Proceedings, volume 4973 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78756-9 (print), 3-540-78757-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78757-0.

Marciniak:2009:ANL

[MM09a] Małgorzata Marciniak and Agnieszka Mykowiecka, editors. Aspects of Natural Language Processing: Essays Dedicated to Leonard Bolc on the Occasion of His 75th Birthday, volume 5070 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04734-3 (print), 3-642-04735-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
04735-0.

Matousek:2009:TSD

[MM09b] Václav Matoušek and Pavel Mautner, editors. Text, Speech and Dialogue: 12th International Conference, TSD 2009, Pilsen, Czech Republic, September 13–17, 2009. Proceedings, volume 5729 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04207-4 (print), 3-642-04208-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04208-9.

Mery:2010:IFM

[MM10] Dominique Méry and Stephan Merz, editors. Integrated Formal Methods: 8th International Conference, IFM 2010, Nancy, France, October 11–14, 2010. Proceedings, volume 6396 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16264-9 (print), 3-642-16265-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16265-7.

Marchand-Maillet:2007:AMR

[MMBND07] Stéphane Marchand-Maillet, Eric Bruno, Andreas Nürnberger, and Marcin Detyniecki, editors. Adaptive Multimedia Retrieval: User, Context, and Feedback: 4th International Workshop, AMR 2006, Geneva, Switzerland, July 27–28, 2006, Revised Selected Papers, volume 4398 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71544-4 (print), 3-540-71545-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71545-0.

Moreno:2005:ESB

[MMC05] J. Manuel Moreno, Jordi Madrenas, and Jordi Cosp, editors. Evolvable Systems: From Biology to Hardware: 6th International Conference, ICES 2005, Sitges, Spain, September 12–14, 2005. Proceedings, volume 3637 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28736-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28736-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3637.

Magedanz:2005:OMI

[MMD05] Thomas Magedanz, Edmundo R. M. Madeira, and Petre (Petre) Dini, editors. Operations and management in IP-based networks: 5th IEEE International Workshop on IP Operations and Management, IPOM 2005, Barcelona, Spain, October 26–28, 2005: proceedings, volume 3751 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29356-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 .I33 2005.

URL http://www.springerlink.
com/content/978-3-540-
29356-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3751.

Masthoff:2012:UMA

[MMDN12] Judith Masthoff, Bamshad Mobasher, Michel C. Desmarais, and Roger Nkambou, editors. User Modeling, Adaptation, and Personalization: 20th International Conference, UMAP 2012, Montreal, Canada, July 16– 20, 2012. Proceedings, volume 7379 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31453-8 (print), 3-642-31454-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31454-4.

Midkiff:2001:LCP

[MMG^{+}01] Samuel P. (Samuel Pratt) Midkiff, J. E. Moreira, M. Gupta, S. Chatterjee, J. Ferrante, J. Prins, W. Pugh, and C.-W. Tseng, editors. Languages and compilers for parallel computing: 13th International Workshop, LCPC 2000, Yorktown Heights, NY, USA, August 10–12, 2000: Revised Papers, volume 2017 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42862-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2017; QA76.58 .W656 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2017.
htm;   http://www.springerlink.
com/content/978-3-540-
42862-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2017.

Meseguer:2010:CTA

[MMG10] Pedro Meseguer, Lawrence Mandow, and Rafael M. Gasca, editors. Current Topics in Artificial Intelligence: 13th Conference of the Spanish Association for Artificial Intelligence, CAEPIA 2009, Seville, Spain, November 9–13, 2009. Selected Papers, volume 5988 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14263-X (print), 3-642-14264-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14264-2.

Mossakowski:2007:ACC

[MMH07] Till Mossakowski, Ugo Montanari, and Magne Haveraaen, editors. Algebra and Coalgebra in Computer Science: Second International Conference, CALCO 2007, Bergen, Norway, August 20–24, 2007. Proceedings, volume 4624 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73857-6 (print), 3-540-73859-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73859-6.

Morvan:1998:SAS

[MMK98] Michel Morvan, Christoph Meinel, and Daniel Krob, editors. STACS 98: 15th annual Symposium on Theoretical Aspects of Computer Science, Paris, France, February 25–27, 1998: Proceedings, volume 1373 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64230-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5.S956 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1373.
htm;   http://www.springerlink.
com/content/978-3-540-
64230-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1373.

Miyano:2005:RCM

[MMK^{+}05] Satoru Miyano, Jill Mesirov, Simon Kasif, Simon Kasif, Sorin Istrail, Pavel Pevzner, and Michael Waterman, editors. Research in Computational Molecular Biology: 9th Annual International Conference, RECOMB 2005 Cambridge, MA, USA, May 14–18, 2005 Proceedings, volume 3500 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25866-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25866-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3500;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b135594.

Mjolsnes:2008:PKI

[MMK08] Stig F. Mjølsnes, Sjouke Mauw, and Sokratis K. Katsikas, editors. Public Key Infrastructure: 5th European PKI Workshop: Theory and Practice, EuroPKI 2008 Trondheim, Norway, June 16–17, 2008 Proceedings, volume 5057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69484-6 (print), 3-540-69485-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69485-4.

Maier:1997:DPS

[MML97] Elisabeth Maier, Marion Mast, and Susann Luperfoy, editors. Dialogue processing in spoken language systems: ECAI ’96 workshop, Budapest, Hungary, August 13, 1996: revised papers, volume 1236 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63175-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 D5 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1236.htm;
 http://www.springerlink.
com/content/978-3-540-
63175-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1236.

Malenfant:2001:OOT

[MMM01] Jacques Malenfant, Sabine Moisan, and Ana M. D. Moreira, editors. Objectoriented technology: ECOOP 2000 workshop reader: ECOOP 2000 workshops, panels, and posters, Sophia Antipolis and Cannes, France, June 12–16, 2000: Proceedings, volume 1964 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41513-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E24 2000a; QA76.9.O35 O27 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1964.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        41513-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1964.

Montoyo:2005:NLP

[MMM05] Andrés Montoyo, Rafael Muñoz, and Elisabeth Métais, editors. Natural Language Processing and Information Systems: 10th International Conference on Applications of Natural Language to Information Systems, NLDB 2005 Alicante, Spain, June 15– 17, 2005 Proceedings, volume 3513 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26031-5 (print), 3-540-32110-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26031-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3513;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136569.

McGuinness:2010:PAD

[MMM10] Deborah L. McGuinness, James R. Michaelis, and Luc Moreau, editors. Provenance and Annotation of Data and Processes: Third International Provenance and Annotation Workshop, IPAW 2010, Troy, NY, USA, June 15–16, 2010. Revised Selected Papers, volume 6378 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17818-9 (print), 3-642-17819-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17819-1.

Munoz:2011:NLP

[MMM11] Rafael Muñoz, Andrés Montoyo, and Elisabeth Métais, editors. Natural Language Processing and Information Systems: 16th International Conference on Applications of Natural Language to Information Systems, NLDB 2011, Alicante, Spain, June 28–30, 2011. Proceedings, volume 6716 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22326-5 (print), 3-642-22327-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22327-3.

Mira:1997:BAC

[MMMD97] J. Mira, Joan Cabestany Moncusi, and Roberto Moreno-Diaz, editors. Biological and artificial computation: from neuroscience to technology: International Work-Conference on Artificial Neural Networks, IWANN 97, Lanzarote, Canary Islands, Spain, June 4–6, 1997: proceedings, volume 1240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63047-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I588 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1240.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63047-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1240.

Miglioli:1996:TPA

[MMMO96] Pierangelo Miglioli, . Moscato, D. Mundici, and M. Ornaghi, editors. Theorem proving with analytic tableaux and related methods: 5th International Workshop, TABLEAUX ’96, Terrasini, Palermo, Italy, May 15–17, 1996: proceedings, volume 1071 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61208-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T33 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1071.htm;
 http://www.springerlink.
com/content/978-3-540-
61208-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1071.

Mahjoub:2012:COS

[MMMP12] A. Ridha Mahjoub, Vangelis Markakis, Ioannis Milis, and Vangelis Th. Paschos, editors. Combinatorial Optimization: Second International Symposium, ISCO 2012, Athens, Greece, April 19–21, 2012, Revised Selected Papers, volume 7422 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32146-1 (print), 3-642-32147-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32147-4.

Mehrotra:2011:MAAa

[MMO^{+}11a] Kishan G. Mehrotra, Chilukuri K. Mohan, Jae C. Oh, Pramod K. Varshney, and Moonis Ali, editors. Modern Approaches in Applied Intelligence: 24th International Conference on Industrial Engineering and Other Applications of Applied Intelligent Systems, IEA/ AIE 2011, Syracuse, NY, USA, June 28 — July 1, 2011, Proceedings, Part I, volume 6703 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21821-0 (print), 3-642-21822-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21822-4.

Mehrotra:2011:MAAb

[MMO^{+}11b] Kishan G. Mehrotra, Chilukuri K. Mohan, Jae C. Oh, Pramod K. Varshney, and Moonis Ali, editors. Modern Approaches in Applied Intelligence: 24th International Conference on Industrial Engineering and Other Applications of Applied Intelligent Systems, IEA/ AIE 2011, Syracuse, NY, USA, June 28 — July 1, 2011, Proceedings, Part II, volume 6704 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21826-1 (print), 3-642-21827-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21827-9.

Matousek:1999:TSD

[MMOS99] Vaclav Matoušek, Pavel Mautner, Jana Ocelíková, and Petr Sojka, editors. Text, speech, and dialogue: Second International Workshop TSD’99, Plzen, Czech Republic, September 13– 17, 1999: Proceedings, volume 1692 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66494-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1692. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1692.htm;
 http://www.springerlink.
com/content/978-3-540-
66494-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1692.

Marik:2003:MAS

[MMP03] Vladimir Marik, Jörg P. Müller, and Michal Pechouek, editors. Multiagent systems and applications III: 3rd International Central and Eastern European Conference on Multi-Agent Systems, CEEMAS 2003, Prague, Czech Republic, June 16–18, 2003: Proceedings, volume 2691 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40450-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.76.I58 C43 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2691.htm;
http://www.springerlink.
com/content/978-3-540-
40450-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2691.

Matousek:2005:TSD

[MMP05] Václav Matoušek, Pavel Mautner, and Tomáš Pavelka, editors. Text, speech and dialogue: 8th international conference, TSD 2005, Karlovy Vary, Czech Republic, September 12–15, 2005: proceedings, volume 3658 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28789-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 T47 2005.

URL http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3658;
http://www.springerlink.
com/content/978-3-540-
28789-6.

Masulli:2007:AFS

[MMP07] Francesco Masulli, Sushmita Mitra, and Gabriella Pasi, editors. Applications of Fuzzy Sets Theory: 7th International Workshop on Fuzzy Logic and Applications, WILF 2007, Camogli, Italy, July 7–10, 2007. Proceedings, volume 4578 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73399-X (print), 3-540-73400-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73400-0.

Mattson:1991:AAA

[MMR91] H. F. Mattson, Teo Mora, and T. R. N. Rao, editors. Applied algebra, algebraic algorithms, and error-correcting codes: 9th International Symposium, AAECC-9, New Orleans, LA, USA, October 7–11, 1991: proceedings, volume 539 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54522-0 (Berlin), 0-387-54522-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0539.htm;
http://www.springerlink.
com/content/978-3-540-
54522-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=539.

MeyeraufderHeide:1993:PAT

[MMR93] F. Meyer auf der Heide, B. Monien, and Arnold Rosenberg, editors. Parallel architectures and their efficient use: First Heinz Nixdorf Symposium, Paderborn, Germany, November 11– 13, 1992: proceedings, volume 678 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56731-3 (Berlin), 0-387-56731-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .H45 1992. DM52.00, US$39.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0678.
htm;   http://www.springerlink.
com/content/978-3-540-
56731-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=678.

Marchiori:2007:ECM

[MMR07] Elena Marchiori, Jason H. Moore, and Jagath C. Rajapakse, editors. Evolutionary computation, machine learning, and data mining in bioinformatics: 5th European Conference, EvoBIO 2007, Valencia, Spain, April 11– 13, 2007: proceedings, volume 4447 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-71782-X (paperback). LCCN QH324.2 .E93 2007.

Mineau:1993:CGK

[MMS93] Guy W. Mineau, B. Moulin, and John F. Sowa, editors. Conceptual graphs for knowledge representation: First International Conference on Conceptual Structures, ICCS’93, Quebec City, Canada, August 4–7, 1993: proceedings, volume 699 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56979-0 (Berlin), 0-387-56979-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387 .I58 1993.

URL http://link.springer-
ny.com/link/service/

series/0558/tocs/t0699.
htm;   http://www.springerlink.
com/content/978-3-540-
56979-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=699.

Masuda:1997:WCA

[MMT97] Takashi Masuda, Y. (Yoshifumi) Masunaga, and Michiharu Tsukamoto, editors. Worldwide computing and its applications: international conference, WWCA ’97, Tsukuba, Japan, March 10–11, 1997: proceedings, volume 1274 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63343-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5.W85 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1274.
htm;   http://www.springerlink.
com/content/978-3-540-
63343-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1274.

Mohamed:2008:TPH

[MMT08] Otmane Ait Mohamed, César Muñoz, and Sofiène Tahar, editors. Theorem Proving in Higher Order Logics: 21st International Conference, TPHOLs 2008, Montreal, Canada, August 18–21, 2008. Proceedings, volume 5170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-71065-5 (print), 3-540-71067-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71067-7.

Miguet:1996:DGC

[MMU96] Serge Miguet, Annick Montanvert, and Stephane Ubeda, editors. Discrete geometry for computer imagery: 6th International Workshop, DGCI ’96, Lyon, France, November 13–15, 1996: proceedings, volume 1176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-62005-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .D525 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1176.
htm;   http://www.springerlink.
com/content/978-3-540-
62005-1;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1176.

Marik:2004:HMA

[MMV04] Vladimír Mařík, Duncan C. McFarlane, and Paul Valckenaers, editors. Holonic and Multi-Agent Systems for Manufacturing: First International Conference on Industrial Applications of Holonic and Multi-Agent Systems, HoloMAS 2003 Prague, Czech Republic, September 1–3, 2003 Proceedings, volume 2744 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-40751-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TS155.63 .I59 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2744.htm;
http://www.springerlink.
com/content/978-3-540-
40751-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2744;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11831.

Maitra:2006:PCI

[MMV06] Subhamoy Maitra, C. E. Veni Madhavan, and Ramarathnam Venkatesan, editors. Progress in Cryptology — INDOCRYPT 2005: 6th International Conference on Cryptology in India, Bangalore, India, December 10–12, 2005. Proceedings, volume 3797 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30805-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30805-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3797.

Mercer:1998:AAI

[MN98] Robert E. Mercer and Eric Neufeld, editors. Advances in artificial intelligence: 12th Biennial Conference of the Canadian Society for Computational Studies of Intelligence, AI’98 Vancouver, BC, Canada, June 18–20, 1998: Proceedings, volume 1418 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64575-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1418. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1418.htm;
 http://www.springerlink.
com/content/978-3-540-
64575-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1418.

Manolopoulos:2002:ADI

[MN02a] Yannis Manolopoulos and Pavol Návrat, editors. Advances in Databases and Information Systems: 6th East European Conference, ADBIS 2002, Bratislava, Slovakia, September 8– 11, 2002. Proceedings, volume 2435 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44138-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 A3396 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2435.htm;
 http://www.springerlink.
com/content/978-3-540-
44138-0;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2435.

Mattern:2002:PCF

[MN02b] Friedemann Mattern and Mahmoud Naghshineh, editors. Pervasive Computing: First International Conference, Pervasive 2002, Zürich, Switzerland, August 26– 28, 2002. Proceedings, volume 2414 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44060-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.59 .P47 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2414.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44060-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2414.

Meyer:2012:ESE

[MN12a] Bertrand Meyer and Martin Nordio, editors. Empirical Software Engineering and Verification: International Summer Schools, LASER 2008-2010, Elba Island, Italy, Revised Tutorial Lectures, volume 7007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25230-3 (print), 3-642-25231-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25231-0.

Meyer:2012:TPS

[MN12b] Bertrand Meyer and Martin Nordio, editors. Tools for Practical Software Verification: LASER, International Summer School 2011, Elba Island, Italy, Revised Tutorial Lectures, volume 7682 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35745-8 (print), 3-642-35746-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35746-6.

Mobasher:2006:AWM

[MNLM06] Bamshad Mobasher, Olfa Nasraoui, Bing Liu, and Brij Masand, editors. Advances in Web Mining and Web Usage Analysis: 6th International Workshop on Knowledge Discovery on the Web, WebKDD 2004, Seattle, WA, USA, August 22–25, 2004, Revised Selected Papers, volume 3932 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47127-8 (print), 3-540-47128-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47128-8.

Medhi:2007:IOM

[MNPW07] Deep Medhi, José Marcos Nogueira, Tom Pfeifer, and S. Felix Wu, editors. IP Operations and Management: 7th IEEE International Workshop, IPOM 2007 San José, USA, October 31 — November 2, 2007 Proceedings, volume 4786 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75852-6 (print), 3-540-75853-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
75853-2.

Mosses:1995:TTP

[MNS95] Peter Mosses, M. Nielsen, and Michael I. Schwartzbach, editors. TAPSOFT ’95: theory and practice of software development: 6th International Joint Conference CAAP/FASE, Aarhus, Denmark, May 22–26, 1995: proceedings, volume 915 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59293-8, 0-387-59293-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I58 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0915.
htm;   http://www.springerlink.
com/content/978-3-540-
59293-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=915.

Malek:2005:SASb

[MNS05] Miroslaw Malek, Edgar Nett, and Neeraj (Neeraj) Suri, editors. Service availability: Second International Service Availability Symposium, ISAS 2005, Berlin, Germany, April 25–26, 2005: revised selected papers, volume 3694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29103-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I5755 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29103-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3694.

Misra:2006:FFM

[MNS06] Jayadev Misra, Tobias Nipkow, and Emil Sekerinski, editors. FM 2006: Formal Methods: 14th International Symposium on Formal Methods, Hamilton, Canada, August 21–27, 2006. Proceedings, volume 4085 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37215-6 (print), 3-540-37216-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37216-5.

Marek:1995:LPN

[MNT95] V. W. Marek, Anil Nerode, and Miroslaw Truszczynski, editors. Logic programming and nonmonotonic reasoning: third international conference, LPNMR ’95, Lexington, KY, USA, June 24–26, 1995: proceedings, volume 928 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59487-6 (Berlin), 0-387-59487-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 L74 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0928.htm;
 http://www.springerlink.
com/content/978-3-540-
59487-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=928.

Murakami:2005:UCS

[MNTY05] Hitomi Murakami, Hideyuki Nakashima, Hideyuki Tokuda, and Michiaki Yasumura, editors. Ubiquitous Computing Systems: Second International Symposium, UCS 2004, Tokyo, Japan, November 8–9, 2004: Revised Selected Papers, volume 3598 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27893-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5915 .U276 2004. URL http://

                                        springerlink.metapress.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3598;   http:
                                        //www.springerlink.
                                        com/content/978-3-540-
                                        27893-1.

Munson:1998:PDD

[MNW98] Ethan V. Munson, Charles K. Nicholas, and Derick Wood, editors. Principles of digital document processing: 4th International Workshop, PODDP ’98, Saint-Malo, France, March 29–30, 1998: Proceedings, volume 1481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65086-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575.P63 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1481.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-

65086-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1481.

Marshall:2001:ANI

[MNW01] Ian W. Marshall, Scott M.

Nettles, and Naoki Wakamiya, editors. Active networks: IFIP-TC6 Third International Working Conference, IWAN, 2001, Philadelphia, PA, USA, September 30–October 2, 2001: Proceedings, volume 2207 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42678-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2207; TK5105.5 .I95 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2207.
htm;   http://www.springerlink.
com/content/978-3-540-
42678-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2207.

Meyer:2008:BAF

[MNW08] Bertrand Meyer, Jerzy R.

Nawrocki, and Bartosz Walter, editors. Balancing Agility and Formalism in Software Engineering: Second IFIP TC 2 Central and East European Conference on Software Engineering Techniques, CEE-SET 2007, Poznan, Poland, October 10–12, 2007, Revised Selected Papers, volume 5082 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85278-6 (print), 3-540-85279-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85279-7.

Mandoiu:2009:BRA

[MNZ09] Ion Măndoiu, Giri Narasimhan, and Yanqing Zhang, editors. Bioinformatics Research and Applications: 5th International Symposium, ISBRA 2009 Fort Lauderdale, FL, USA, May 13–16, 2009 Proceedings, volume 5542 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01550-6 (print), 3-642-01551-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
01551-9.

Monien:1992:DSE

[MO92a] B. Monien and Thomas Ottmann, editors. Data structures and efficient algorithms: final report on the DFG special joint initiative, volume 594 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55488-2 (Berlin), 0-387-55488-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 D36 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0594.htm;
http://www.springerlink.
com/content/978-3-540-
55488-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=594.

Papers presented at a symposium held at the Freie Universität Berlin, Sept. 23–26, 1991.

Myers:1992:CCS

[MO92b] J. P. Myers and Michael J.

O’Donnell, editors. Constructivity in computer science: summer symposium, San Antonio, TX, June 19–22, 1991, proceedings, volume 613 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55631-1 (Berlin), 0-387-55631-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 C655 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0613.htm;
 http://www.springerlink.
com/content/978-3-540-
55631-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=613.

Contains papers presented at the Symposium on Constructivity in Computer Science held in San Antonio, Tex., June 19–22, 1991, sponsored by Trinity University, the University of Chicago, and the Association for Symbolic Logic.

Muller-Olm:1997:MCV

[MO97] Markus Muller-Olm. Modular compiler verification: a refinementalgebraic approach advocating stepwise abstraction, volume 1283 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63406-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 250 pp. LCCN QA76.76.C65M85 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1283.htm;
 http://www.springerlink.
com/content/978-3-540-
63406-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1283.

Malenfant:2005:OOT

[MØ05a] Jacques Malenfant and Bjarte M. Østvold, editors. Object-oriented technology: ECOOP 2004 workshop reader: ECOOP 2004 workshops, Oslo, Norway, June 14– 18, 2004: final reports, volume 3344 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-23988-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35 E33 2004b. URL http:

//www.springerlink.
com/content/978-3-540-
23988-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3344;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b104146.

Morrison:2005:SAE

[MO05b] Ron Morrison and Flavio Oquendo, editors. Software Architecture: 2nd European Workshop, EWSA 2005, Pisa, Italy, June 13–14, 2005. Proceedings, volume 3527 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26275-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .E97 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26275-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3527.

Muller-Olm:2006:VCF

[MO06] Markus Müller-Olm, editor. Variations on Constants: Flow Analysis of Sequential and Parallel Programs, volume 3800 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-45385-7 (print), 3-540-45386-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 171 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45386-4.

Medina:2008:FCA

[MO08] Raoul Medina and Sergei Obiedkov, editors. Formal Concept Analysis: 6th International Conference, ICFCA 2008, Montreal, Canada, February 25–28, 2008. Proceedings, volume 4933 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78136-6 (print), 3-540-78137-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78137-0.

Marin:2006:CTA

[MOBS06] Roque Marín, Eva Onaindía, Alberto Bugarín, and José Santos, editors. Current Topics in Artificial Intelligence: 11th Conference of the Spanish Association for Artificial Intelligence, CAEPIA 2005, Santiago de Compostela, Spain, November 16–18, 2005, Revised Selected Papers, volume 4177 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45914-6 (print), 3-540-45915-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        45915-6.

Marti:2010:DMI

[MOFM10] Joan Martí, Arnau Oliver, Jordi Freixenet, and Robert Martí, editors. Digital Mammography: 10th International Workshop, IWDM 2010, Girona, Catalonia, Spain, June 16–18, 2010. Proceedings, volume 6136 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13665-6 (print), 3-642-13666-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13666-5.

Mohring:1991:GTC

[Moh91] R. H. Mohring, editor. Graph-theoretic concepts in computer science: 16th International Workshop WG’90, Berlin, Germany, June 20–22, 1990: proceedings, volume 484 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-53832-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G7172 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0484.htm;
 http://www.springerlink.
com/content/978-0-387-
53832-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=484.

The 16th International Workshop on Graph-Theoretic Concepts in Computer Science (WG’90) organized by the group “Discrete Mathematics and Computer Science” of the Department of Mathematics of the Technical University of Berlin.

Mohring:1997:GTC

[Moh97] Rolf H. Mohring, editor.

Graph-theoretic concepts in computer science: 23rd International Workshop, WG’97, Berlin, Germany, June 18–20, 1997, proceedings, volume 1335 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63757-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1335.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1335.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63757-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1335.

Ma:2012:SGD

[MOH^{+}12] Minhua Ma, Manuel Fradinho Oliveira, Jannicke Baalsrud Hauge, Heiko Duin, and Klaus-Dieter Thoben, editors. Serious Games Development and Applications: Third International Conference, SGDA 2012, Bremen, Germany, September 26– 29, 2012. Proceedings, volume 7528 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33686-8 (print), 3-642-33687-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33687-4.

Moller:1995:MPC

[Mol95] Bernhard Moller, editor. Mathematics of program construction: third international conference, MPC ’95, Kloster Irsee, Germany, July 17–21, 1995: proceedings, volume 947 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60117-1 (Berlin: softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5535 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0947.htm;
 http://www.springerlink.
com/content/978-3-540-
60117-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=947.

Montangero:1996:SPT

[Mon96] Carlo Montangero, editor. Software process technology: 5th European workshop, EWSPT ’96, Nancy, France, October 9–11, 1996: proceedings, volume 1149 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61771-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47E87 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1149.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61771-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1149.

Montello:2001:SIT

[Mon01] Daniel R. Montello, editor. Spatial information theory: foundations of geographic information science: International Conference, COSIT 2001, Morro Bay, CA, USA, September 19–23, 2001: Proceedings, volume 2205 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42613-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2205; G70.212 .C68 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2205.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42613-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2205.

Moortgat:2001:LAC

[Moo01] Michael Moortgat, editor. Logical aspects of computational linguistics: Third International Conference, LALC ’98, Grenoble, France, December 14– 16, 1998: Selected Papers, volume 2014 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2014; P98.5.L63 L33 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2014.
htm;   http://www.springerlink.
com/content/978-3-540-
42251-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2014.

Macdonald:2008:AIR

[MOP^{+}08] Craig Macdonald, Iadh Ounis, Vassilis Plachouras, Ian Ruthven, and Ryen W. White, editors. Advances in Information Retrieval: 30th European Conference on IR Research, ECIR 2008, Glasgow, UK, March 30– April 3, 2008. Proceedings, volume 4956 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78645-7 (print), 3-540-78646-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78646-7.

Ma:2011:SGD

[MOP11] Minhua Ma, Manuel Fradinho Oliveira, and João Madeiras Pereira, editors. Serious Games Development and Applications: Second International Conference, SGDA 2011, Lisbon, Portugal, September 19– 20, 2011. Proceedings, volume 6944 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23833-5 (print), 3-642-23834-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23834-5.

Mora:1989:AAA

[Mor89a] Teo Mora, editor. Applied algebra, algebraic algorithms, and errorcorrecting codes: 6th international conference, AAECC-6, Rome, Italy, July 4–8, 1988: proceedings, volume 357 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51083-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 1988. US$36.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0357.
htm;   http://www.springerlink.
com/content/978-0-387-
51083-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=357.

Morik:1989:KRO

[Mor89b] Katharina Morik. Knowledge representation and organization in machine learning, volume 347 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-50768-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 319 pp. LCCN Q325 .M671 1989. DM45.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0347.
htm;   http://www.springerlink.
com/content/978-0-387-
50768-2;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=347.

Moran:1995:AAL

[Mor95] F. Moran, editor. Advances in artificial life: Third European Conference on Artificial Life, Granada, Spain, June 4–6, 1995: proceedings, volume 929 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59496-5 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .E87 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0929.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        59496-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=929.

Morisio:2006:RSC

[Mor06] Maurizio Morisio, editor. Reuse of Off-the-Shelf Components: 9th International Conference on Software Reuse, ICSR 2006 Turin, Italy, June 12–15, 2006 Proceedings, volume 4039 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34606-6 (print), 3-540-34607-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34607-4.

Mossenbock:1997:MPL

[Mos97] Hanspeter Mossenbock, editor. Modular programming languages: Joint Modular Languages Conference, JML ’97, Linz, Austria, March 19–21, 1997: proceedings, volume 1204 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62599-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1204.htm;
http://www.springerlink.
com/content/978-3-540-
62599-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1204.

Moskowitz:2001:IHI

[Mos01] Ira S. Moskowitz, editor.

Information hiding: 4th International Workshop, IH 2001, Pittsburgh, PA, USA, April 25–27, 2001: Proceedings, volume 2137 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42733-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I5414 2001; QA76.9.A25 I49 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2137.htm;
 http://www.springerlink.
com/content/978-3-540-
42733-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2137.

Mosses:2004:CRM

[Mos04] Peter D. Mosses, editor. CASL Reference Manual: The Complete Documentation of the Common Algebraic Specification Language, volume 2960 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21301-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2960.htm;
 http://www.springerlink.

com/content/978-3-540-
21301-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2960;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96103.

Modgil:2012:TAF

[MOT12] Sanjay Modgil, Nir Oren, and Francesca Toni, editors. Theorie and Applications of Formal Argumentation: First International Workshop, TAFA 2011. Barcelona, Spain, July 16–17, 2011, Revised Selected Papers, volume 7132 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29183-X (print), 3-642-29184-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29184-5.

Muggleton:2007:ILP

[MOTN07] Stephen Muggleton, Ramon Otero, and Alireza Tamaddoni-Nezhad, editors. Inductive Logic Programming: 16th International Conference, ILP 2006, Santiago de Compostela, Spain, August 24–27, 2006, Revised Selected Papers, volume 4455 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73846-0 (print), 3-540-73847-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73847-3.

Milne:1993:CHD

[MP93] George J. Milne and Laurence Pierre, editors. Correct hardware design and verification methods: IFIP WG10.2 Advanced Research Working Conference, CHARME ’93, Arles, Frances [sic], May 24–26, 1993: proceedings, volume 683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56778-X (Berlin), 3-540-56778-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874 .A3353 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0683.htm;
 http://www.springerlink.
com/content/978-0-387-

56778-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=683.

“Papers presented at the ‘Advanced Research Working Conference on Correct HARdware Design MEthodologies’ held in Arles (France)”–Foreword. “Organized by the ES-PRIT Working Group 6018 ‘CHARME-2’ and the Université de Provence (Marseille), in cooperation with IFIP WG10.2”– Foreword.

Masuch:1994:KRR

[MP94] Michael Masuch and Laszlo Polos. Knowledge representation and reasoning under uncertainty: logic at work, volume 808 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58095-6 (New York), 3-540-58095-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). 237 pp. LCCN Q387 .K563 1994. DM52.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0808.
htm;   http://www.springerlink.
com/content/978-0-387-
58095-1;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=808.

Mayr:1995:SAS

[MP95a] Ernst Mayr and Claude Puech, editors. STACS 95: 12th Annual Symposium on Theoretical Aspects of Computer Science, Munich, Germany, March 2–4, 1995: proceedings, volume 900 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59042-0 (Berlin: softcover), 0-387-59042-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S956 1995. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0900.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        59042-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=900.

Muller:1995:CSC

[MP95b] Silvia M. Muller and Wolfgang J. Paul. The complexity of simple computer architectures, volume 995 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60580-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 270 pp. LCCN QA76.9.A73 M84 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0995.
htm;   http://www.springerlink.
com/content/978-3-540-
60580-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=995.

McCune:1996:ADE

[MP96] W. (William) McCune and R. (Ranganathan) Padmanabhan. Automated deduction in equational logic and cubic curves, volume 1095 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61398-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 231 pp. LCCN QA76.9.A96M38 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1095.
htm;   http://www.springerlink.
com/content/978-3-540-
61398-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1095.

Maher:1998:PPC

[MP98] Michael Maher and Jean-François Puget, editors. Principles and Practice of Constraint Programming— CP98: 4th International Conference, CP98, Pisa, Italy, October 26– 30, 1998. Proceedings, volume 1520 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65224-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612.C7 1998. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1520.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65224-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1520.

Mira:2001:PSIa

[MP01a] José Mira and Alberto Prieto, editors. Proceedings: Sixth International Work-Conference on Artificial and Natural Neural Networks, IWANN 2001, Granada, Spain, June 13–15, 2001, Part I. Connectionist Models of Neurons, Learning Processes, and Artificial Intelligence, volume 2084 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42235-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2084-2085; QP363.3 .I586 2001 pt.1 (2001); QP363.3 .I586 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2084.htm;
 http://www.springerlink.
com/content/978-3-540-
42235-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2084.

Contents: pt. 1. Connectionist models of neurons, learning processes, and artificial intelligence

— pt. 2. Bio-inspired applications of connectionism.

Mira:2001:PSIb

[MP01b] José Mira and Alberto Prieto, editors. Proceedings: Sixth International Work-Conference on Artificial and Natural Neural Networks, IWANN 2001, Granada, Spain, June 13–15, 2001, Part II. Bio-Inspired Applications of Connectionism, volume 2085 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42237-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2084-2085; QP363.3 .I586 2001 pt.1 (2001); QP363.3 .I586 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2084.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42237-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2085.

Contents: pt. 1. Connectionist models of neurons, learning processes, and artificial intelligence

— pt. 2. Bio-inspired applications of connectionism.

Maler:2003:HSC

[MP03a] O. Maler and A. Pnueli, editors. Hybrid systems: computation and control: 6th International Workshop, HSCC 2003, Prague, Czech Republic, April 3–5, 2003: Proceedings, volume 2623 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00913-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.38 .H73 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2623.htm;
 http://www.springerlink.
com/content/978-3-540-
00913-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2623.

Mayer:2003:ARA

[MP03b] Marta Cialdea Mayer and Fiora Pirri, editors. Automated Reasoning with Analytic Tableaux and Related Methods: International Conference, TABLEAUX 2003, Rome, Italy, September 9–12, 2003: Proceedings, volume 2796 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40787-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T33 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2796.
htm;   http://www.springerlink.
com/content/978-3-540-
40787-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2796;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b12030.

Maicher:2006:CTM

[MP06] Lutz Maicher and Jack Park, editors. Charting the Topic Maps Research and Applications Landscape: First International Workshop on Topic Map Research and Applications, TMRA 2005, Leipzig, Germany, October 6–7, 2005, Revised Selected Papers, volume 3873 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32527-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32527-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3873.

Mavronicolas:2009:AGT

[MP09] Marios Mavronicolas and Vicky G. Papadopoulou, editors. Algorithmic Game Theory: Second International Symposium, SAGT 2009, Paphos, Cyprus, October 18–20, 2009. Proceedings, volume 5814 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04644-4 (print), 3-642-04645-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04645-2.

Manna:2010:TVE

[MP10a] Zohar Manna and Doron A.

Peled, editors. Time for Verification: Essays in Memory of Amir Pnueli, volume 6200 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13753-9 (print), 3-642-13754-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13754-9.

Martinelli:2010:PKI

[MP10b] Fabio Martinelli and Bart Preneel, editors. Public Key Infrastructures, Services and Applications: 6th European Workshop, EuroPKI 2009, Pisa, Italy, September 10–11, 2009, Revised Selected Papers, volume 6391 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16440-4 (print), 3-642-16441-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16441-5.

Modersheim:2012:TSA

[MP12] Sebastian Mödersheim and Catuscia Palamidessi, editors. Theory of Security and Applications: Joint Workshop, TOSCA 2011, Saarbrücken, Germany, March 31 — April 1, 2011, Revised Selected Papers, volume 6993 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27374-2 (print), 3-642-27375-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27375-9.

Marik:2005:HMA

[MPB05] Vladimír Marík, Michal Pechoucek, and Robert W. Brennan, editors. Holonic and multi-agent systems for manufacturing: Second International Conference on Industrial Applications of Holonic and Multi-Agent Systems, HoloMAS 2005, Copenhagen, Denmark, August 22–24, 2005: proceedings, volume 3593 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28237-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28237-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3593.

Marchiori:2007:WRR

[MPdSM07] Massimo Marchiori, Jeff Z.

Pan, and Christian de Sainte Marie, editors. Web Reasoning and Rule Systems: First International Conference, RR 2007, Innsbruck , Austria, June 7–8, 2007. Proceedings, volume 4524 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72981-X (print), 3-540-72982-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72982-2.

Mamede:1995:PAI

[MPF95] Nuno J. Mamede and Carlos Pinto-Ferreira, editors. Progress in artificial intelligence: 7th Portuguese Conference on Artificial Intelligence, EPIA ’95, Funchal, Madeira Island, Portugal, October 3–6, 1995: proceedings, volume 990 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60428-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P67 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0990.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60428-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=990.

Macii:2004:ICS

[MPK04] Enrico Macii, Vassilis Paliouras, and Odysseas Koufopavlou, editors. Integrated Circuit and System Design: Power and Timing Modeling, Optimization and Simulation, 14th International Workshop, PATMOS 2004, Santorini, Greece, September 15–17, 2004, Proceedings, volume 3254 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23095-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3254.htm;
 http://www.springerlink.
com/content/978-3-540-
23095-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3254;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100662.

MacNish:1994:LAI

[MPP94] Craig MacNish, David A.

Pearce, and Luis Moniz Pereira, editors. Logics in artificial intelligence: European Workshop JELIA ’94, York, UK, September 5–8, 1994: proceedings, volume 838 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58332-7 (Berlin), 0-387-58332-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .E985 1994.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0838.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        58332-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=838.

Mahonen:2008:TSB

[MPP08] Petri Mähönen, Klaus Pohl, and Thierry Priol, editors. Towards a Service-Based Internet: First European Conference, ServiceWave 2008, Madrid, Spain, December 10–13, 2008. Proceedings, volume 5377 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89896-4 (print), 3-540-89897-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89897-9.

Mauri:2005:MCI

[MPPJ^{+}05] Giancarlo Mauri, Gheorghe Păun, Mario J.

Pérez-Jiménez, Grzegorz Rozenberg, and Arto Salomaa, editors. Membrane computing: 5th international workshop, WMC 2004: Milan, Italy, June 14–16, 2004: Revised selected and invited papers, volume 3365 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25080-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.887 .W43 2004. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3365;
 http://www.springerlink.
com/content/978-3-540-
25080-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3365;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106721.

Min:2011:STE

[MPPU11] Sang Lyul Min, Robert Pettit, Peter Puschner, and Theo Ungerer, editors. Software Technologies for Embedded and Ubiquitous Systems: 8th IFIP WG 10.2 International Workshop, SEUS 2010, Waidhofen/ Ybbs, Austria, October 13–15, 2010. Proceedings, volume 6399 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16255-X (print), 3-642-16256-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16256-5.

Maudet:2007:AMA

[MPR07] Nicolas Maudet, Simon Parsons, and Iyad Rahwan, editors. Argumentation in Multi-Agent Systems: Third International Workshop, ArgMAS 2006 Hakodate, Japan, May 8, 2006 Revised Selected and Invited Papers, volume 4766 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75525-X (print), 3-540-75526-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75526-5.

McBurney:2012:AMA

[MPR12] Peter McBurney, Simon Parsons, and Iyad Rahwan, editors. Argumentation in Multi-Agent Systems: 8th International Workshop, ArgMAS 2011, Taipei, Taiwan, May 3, 2011, Revised Selected Papers, volume 7543 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33151-3 (print), 3-642-33152-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33152-7.

Moller:1993:FPD

[MPS93] Bernhard Moller, Helmut A. Partsch, and Steve Schuman. Formal program development: IFIP TC2/WG 2.1 state-of-the-art report, volume 755 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57499-9. ISSN 0302-9743 (print), 1611-3349 (electronic). 371 pp. LCCN QA76.76.D47 F68 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0755.htm;
 http://www.springerlink.

                                        com/content/978-0-387-
                                        57499-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=755.

Mayr:1998:LPV

[MPS98] Ernst W. Mayr, H. J. (Hans Jürgen) Prömel, and Angelika Steger, editors. Lectures on Proof Verification and Approximation Algorithms, volume 1367 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64201-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 344 pp. LCCN QA76.9.A96L43 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1367.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64201-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1367.

Mullen:2004:FFA

[MPS04] Gary L. Mullen, Alain Poli, and Henning Stichtenoth, editors. Finite Fields and Applications: 7th International Conference, Fq7 Toulouse, France, May 5–9, 2003: Revised Papers, volume 2948 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21324-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA247.3 .I57 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2948.htm;
 http://www.springerlink.
com/content/978-3-540-
21324-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2948;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95905.

Manolopoulos:2006:ADI

[MPS06] Yannis Manolopoulos, Jaroslav Pokorný, and Timos K. Sellis, editors. Advances in Databases and Information Systems: 10th East European Conference, AD-BIS 2006, Thessaloniki, Greece, September 3– 7, 2006. Proceedings, volume 4152 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37899-5 (print), 3-540-37900-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37900-3.

Masulli:2010:CIM

[MPT10] Francesco Masulli, Leif E. Peterson, and Roberto Tagliaferri, editors. Computational Intelligence Methods for Bioinformatics and Biostatistics: 6th International Meeting, CIBB 2009, Genoa, Italy, October 15–17, 2009, Revised Selected Papers, volume 6160 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14570-1 (print), 3-642-14571-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14571-1.

Monien:2001:HPC

[MPV01] Burkhard Monien, Viktor K. Prasanna, and Sriram Vajapeyam, editors. High Performance Computing—HiPC 2001: 8th International Conference, Hyderabad, India, December, 17–20, 2001. Proceedings, volume 2228 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43009-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I5847 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2228.htm;
http://www.springerlink.
com/content/978-3-540-
43009-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2228.

Maglogiannis:2012:AIT

[MPV12] Ilias Maglogiannis, Vassilis Plagianakos, and Ioannis Vlahavas, editors. Artificial Intelligence: Theories and Applications: 7th Hellenic Conference on AI, SETN 2012, Lamia, Greece, May 28–31, 2012. Proceedings, volume 7297 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30447-8 (print), 3-642-30448-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30448-4.

McIlraith:2004:SWI

[MPvH04] Sheila A. McIlraith, Dimitris Plexousakis, and Frank van Harmelen, editors. The Semantic Web—ISWC 2004: Third International Semantic Web Conference, Hiroshima, Japan, November 7–11, 2004, Proceedings, volume 3298 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23798-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3298.htm;
 http://www.springerlink.
com/content/978-3-540-
23798-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3298;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b102467.

Martins:1991:TMS

[MR91] Joao P. Martins and M. Reinfrank, editors. Truth maintenance systems: ECAI-90 workshop, Stockholm, Sweden, August 6, 1990: proceedings, volume 515 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54305-8 (Berlin), 0-387-54305-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q338 .T78 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0515.
htm;   http://www.springerlink.
com/content/978-3-540-
54305-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=515.

“[Papers from] the Truth Maintenance Workshop … held on August 6, 1990, during the European Conference on Artificial Intelligence in Stockholm, Sweden”– Pref.

Montanari:1995:PPC

[MR95] U. Montanari and Francesca Rossi, editors. Principles and practice of constraint programming — CP ’95: first international conference, CP ’95, Cassis, France, September 19–22, 1995: proceedings, volume 976 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60299-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C7 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0976.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60299-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=976.

Marie:1997:CPE

[MR97a] Raymond Marie and Gerardo Rubino, editors. Computer performance evaluation: modelling techniques and tools: 9th international conference, St. Malo, France, June 3–6, 1997: proceedings, volume 1245 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63101-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.E94 I55 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1245.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63101-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1245.

Moggi:1997:CTC

[MR97b] Eugenio Moggi and G. (Giuseppe) Rosolini, editors. Category theory and computer science: 7th international conference, CTCS 97, S. Margheria Ligure, Italy, September 1997: proceedings, volume 1290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63455-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35B53 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1290.
htm;   http://www.springerlink.
com/content/978-3-540-
63455-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1290.

Margenstern:2001:MCU

[MR01] Maurice Margenstern and Yurii Rogozhin, editors. Machines, Computations, and Universality: Third International Conference, MCU 2001: Chisinau, Moldova, May 23–27, 2001: Proceedings, volume 2055 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42121-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .M33 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2055.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42121-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2055.

Mohring:2002:AEA

[MR02] R. H. Möhring and Rajeev Raman, editors. Algorithms — ESA 2002: 10th annual European symposium, Rome, Italy, September 17–21, 2002: Proceedings, volume 2461 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44180-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2461.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44180-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2461.

Mery:2007:AIV

[MR07a] Domingo Mery and Luis Rueda, editors. Advances in Image and Video Technology: Second Pacific Rim Symposium, PSIVT 2007 Santiago, Chile, December 17–19, 2007 Proceedings, volume 4872 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77128-X (print), 3-540-77129-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77129-6.

Miguel:2007:ARA

[MR07b] Ian Miguel and Wheeler Ruml, editors. Abstraction, Reformulation, and Approximation: 7th International Symposium, SARA 2007, Whistler, Canada, July 18–21, 2007. Proceedings, volume 4612 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73579-8 (print), 3-540-73580-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73580-9.

Meseguer:2008:AMS

[MR08] José Meseguer and Grigore Roşu, editors. Algebraic Methodology and Software Technology: 12th International Conference, AMAST 2008 Urbana, IL, USA, July 28–31, 2008 Proceedings, volume 5140 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79979-6 (print), 3-540-79980-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79980-1.

Mouratidis:2011:AIS

[MR11] Haralambos Mouratidis and Colette Rolland, editors. Advanced Information Systems Engineering: 23rd International Conference, CAiSE 2011, London, UK, June 20–24, 2011. Proceedings, volume 6741 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21639-0 (print), 3-642-21640-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21640-4.

Marx:2012:PEC

[MR12a] Dániel Marx and Peter Rossmanith, editors. Parameterized and Exact Computation: 6th International Symposium, IPEC 2011, Saarbrücken, Germany, September 6–8, 2011. Revised Selected Papers, volume 7112 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28049-8 (print), 3-642-28050-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28050-4.

Moot:2012:LCG

[MR12b] Richard Moot and Christian Retoré, editors. The Logic of Categorial Grammars: A Deductive Account of Natural Language Syntax and Semantics, volume 6850 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-31554-2 (print), 3-642-31555-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 297 (est.) pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31555-8.

Moreira:2012:IAA

[MR12c] Nelma Moreira and Rogério Reis, editors. Implementation and Application of Automata: 17th International Conference, CIAA 2012, Porto, Portugal, July 17–20, 2012. Proceedings, volume 7381 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31605-0 (print), 3-642-31606-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31606-7.

Machado:2012:EBI

[MRC12] Penousal Machado, Juan Romero, and Adrian Carballal, editors. Evolutionary and Biologically Inspired Music, Sound, Art and Design: First International Conference, EvoMUSART 2012, Málaga, Spain, April 11–13, 2012. Proceedings, volume 7247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29141-4 (print), 3-642-29142-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29142-5.

Malek:2005:SAFa

[MRK05] Miroslaw Malek, Manfred Reitenspieß, and Jörg Kaiser, editors. Service availability: first International Service Availability Symposium, ISAS 2004, Munich, Germany, May 13–14, 2004: Revised selected papers, volume 3335 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24420-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .S4223 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3335;
 http://www.springerlink.
com/content/978-3-540-
24420-2;   http://www.
springerlink.com/link.
asp?id=rwxx0hrf71xf;

                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3334;   http:
                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105309.

Ma:2007:ECI

[MRN07] Lizhuang Ma, Matthias Rauterberg, and Ryohei Nakatsu, editors. Entertainment Computing

— ICEC 2007: 6th International Conference, Shanghai, China, September 15–17, 2007. Proceedings, volume 4740 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74872-5 (print), 3-540-74873-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74873-1.

McBurney:2011:AMA

[MRP11] Peter McBurney, Iyad Rahwan, and Simon Parsons, editors. Argumentation in Multi-Agent Systems: 7th International Workshop, ArgMAS 2010 Toronto, ON, Canada, May 10, 2010 Revised, Selected and Invited Papers, volume 6614 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21939-X (print), 3-642-21940-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21940-5.

McBurney:2010:AMA

[MRPM10] Peter McBurney, Iyad Rahwan, Simon Parsons, and Nicolas Maudet, editors. Argumentation in Multi-Agent Systems: 6th International Workshop, ArgMAS 2009, Budapest, Hungary, May 12, 2009. Revised Selected and Invited Papers, volume 6057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12804-1 (print), 3-642-12805-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12805-9.

Mycielski:1997:SLC

[MRS97] Jan Mycielski, Grzegorz Rozenberg, and Arto Salomaa. Structures in logic and computer science: a selection of essays, volume 1261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63246-8. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 370 pp. LCCN QA76.S844 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1261.htm;
 http://www.springerlink.
com/content/978-3-540-
63246-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1261.

Marik:2003:DES

[MRŠ03] Vladimír Mařïk, Werner Retschitzegger, and Olga Štĕpánková, editors. Database and Expert Systems Applications: 14th International Conference, DEXA 2003 Prague, Czech Republic, September 1–5, 2003 Proceedings, volume 2736 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40806-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D476 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2736.
htm;   http://www.springerlink.
com/content/978-3-540-
40806-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2736;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11824.

Mele:2007:ABV

[MRSV07] Francesco Mele, Giuliana Ramella, Silvia Santillo, and Francesco Ventriglia, editors. Advances in Brain, Vision, and Artificial Intelligence: Second International Symposium, BVAI 2007, Naples, Italy, October 10–12, 2007. Proceedings, volume 4729 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75554-3 (print), 3-540-75555-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75555-5.

Marquez:2012:CGX

[MRU12] Alberto Márquez, Pedro Ramos, and Jorge Urrutia, editors. Computational Geometry: XIV Spanish Meeting on Computational Geometry, EGC 2011, Dedicated to Ferran Hurtado on the Occasion of His 60th Birthday, Alcalá de Henares, Spain, June 27–30, 2011, Revised Selected Papers, volume 7579 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34190-X (print), 3-642-34191-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34191-5.

Malek:2007:SAI

[MRvM07] Miroslaw Malek, Manfred Reitenspieß, and Aad van Moorsel, editors. Service Availability: 4th International Service Availability Symposium, ISAS 2007, Durham, NH, USA, May 21– 22, 2007. Proceedings, volume 4526 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72735-3 (print), 3-540-72736-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72736-1.

Maximilien:2011:SOC

[MRY^{+}11] E. Michael Maximilien, Gustavo Rossi, Soe-Tsyr Yuan, Heiko Ludwig, and Marcelo Fantinato, editors. Service-Oriented Computing: ICSOC 2010 International Workshops, PAASC, WESOA, SEE, and SOC-LOG, San Francisco, CA, USA, December 7–10, 2010, Revised Selected Papers, volume 6568 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19393-5 (print), 3-642-19394-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19394-1.

Massacci:2009:ESS

[MRZ09] Fabio Massacci, Samuel T.

Redwine Jr., and Nicola Zannone, editors. Engineering Secure Software and Systems: First International Symposium ESSoS 2009, Leuven, Belgium, February 4– 6, 2009. Proceedings, volume 5429 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00198-X (print), 3-642-00199-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00199-4.

Mira:1995:NAN

[MS95] J. Mira and Francisco Sandoval, editors. From natural to artifical neural computation: International Workshop on Artificial Neural Networks, Malaga-Torremolinos, Spain, June 7–9, 1995: proceedings, volume 930 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59497-3 (Berlin), 0-387-59497-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I588 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0930.htm;
 http://www.springerlink.
com/content/978-3-540-
59497-0;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=930.

Margaria:1996:TAC

[MS96a] Tiziana Margaria and Bernhard Steffen, editors. Tools and algorithms for the construction and analysis of systems: second International Workshop, TACAS ’96, Passau, Germany, March 27– 29, 1996: proceedings, volume 1055 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61042-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T33 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1055.htm;
http://www.springerlink.
com/content/978-3-540-
61042-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1055.

McRobbie:1996:ADC

[MS96b] M. A. (Michael A.)

McRobbie and J. K. (John K.) Slaney, editors. Automated deduction, Cade-13: 13th International Conference on Automated Deduction, New Brunswick, NJ, USA, July 30– August 3, 1996: proceedings, volume 1104 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61511-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I57 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1104.
htm;   http://www.springerlink.
com/content/978-3-540-
61511-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1104.

Montanari:1996:CCT

[MS96c] U. (Ugo) Montanari and Vladimiro Sassone, editors. CONCUR ’96: concurrency theory: 7th International Conference, Pisa, Italy, August 26– 29, 1996: proceedings, volume 1119 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61604-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 1996. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t1119.htm;
http://www.springerlink.
com/content/978-3-540-
61604-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1119.

Meyer:1999:FMA

[MS99a] John-Jules Ch. Meyer and Pierre-Yves Schobbens, editors. Formal models of agents: ESPRIT Project ModelAge final workshop Selected Papers, volume 1760 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-67027-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 F67 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1760.htm;
http://www.springerlink.
com/content/978-3-540-
67027-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1760.

Middeldorp:1999:FLP

[MS99b] Aart Middeldorp and Taisuke Sato, editors. Functional and logic programming: 4th Fuji International Symposium, FLOPS’99, Tsukuba, Japan, November 11– 13, 1999: Proceedings, volume 1722 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66677-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1722. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1722.htm;
 http://www.springerlink.
com/content/978-3-540-
66677-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1722.

Masand:2000:WUA

[MS00a] Brij Masand and Myra Spiliopoulou, editors. Web usage analysis and user profiling: international WEBKDD’99 workshop, San Diego, CA, USA, August 15, 1999: Revised Papers, volume 1836 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67818-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 I479 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1836.htm;
 http://www.springerlink.
com/content/978-3-540-
67818-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1836.

Mizoguchi:2000:PTA

[MS00b] Riichiro Mizoguchi and J. K. Slaney, editors. PRICAI 2000, topics in artificial intelligence: 6th Pacific Rim International Conference on Artificial Intelligence, Melbourne, Australia, August 28–September 1, 2000: Proceedings, volume 1886 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67925-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P33 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1886.htm;
 http://www.springerlink.
com/content/978-3-540-
67925-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1886.

Monard:2000:AAI

[MS00c] Maria Carolina Monard and Jaime S. Sichman, editors. Advances in Artificial Intelligence: International Joint Conference, 7th Ibero–American Conference on AI, 15th Brazilian Symposium on AI, IBERAMIA-SBIA 2000, Atibaia, SP, Brazil, November 19–22, 2000. Proceedings, volume 1952 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41276-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1952. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1952.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41276-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1952.

McKay:2002:AAA

[MS02a] Bob McKay and J. K. Slaney, editors. AI 2002: advances in artificial intelligence: 15th Australian Joint Conference on Artificial Intelligence, Canberra, Australia, December 2–6, 2002: Proceedings, volume 2557 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00197-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A97 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2557.htm;
 http://www.springerlink.
com/content/978-3-540-
00197-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2557.

Menezes:2002:PCI

[MS02b] A. J. (Alfred J.) Menezes and Palash Sarkar, editors. Progress in cryptology: INDOCRYPT 2002: Third International Conference on Cryptology in India, Hyderabad, India, December 16–18, 2002: Proceedings, volume 2551 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00263-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I5535 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2551.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00263-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2551.

Also available via the

World Wide Web.

Mount:2002:AEE

[MS02c] David M. Mount and Clifford Stein, editors. Algorithm Engineering and Experiments: 4th International Workshop, ALENEX 2002, San Francicsco, CA, USA, January 4–5, 2002. Revised Papers, volume 2409 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43977-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A42 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2409.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43977-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2409.

Marchesi:2003:EPA

[MS03a] Michele Marchesi and Giancarlo Succi, editors. Extreme programming and agile processes in software engineering: 4th International Conference, XP 2003, Genova, Italy, May 25–29, 2003: Proceedings, volume 2675 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40215-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 X62 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2675.htm;
http://www.springerlink.
com/content/978-3-540-
40215-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2675.

Matwin:2003:ILP

[MS03b] Stan Matwin and Claude Sammut, editors. Inductive logic programming: 12th International Conference, ILP 2002, Sydney, Australia, July 2002: Revised Papers, volume 2583 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00567-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I52 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2583.htm;
 http://www.springerlink.
com/content/978-3-540-
00567-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2583.

Also available via the World Wide Web.

Mendelson:2003:ALM

[MS03c] Shahar Mendelson and Alexander J. Smola, editors. Advanced lectures on machine learning: Machine Learning Summer School 2002, Canberra, Australia, February 11–22, 2002: revised lectures, volume 2600 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00529-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .M344 2002. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t2600.htm;
http://www.springerlink.
com/content/978-3-540-
00529-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2600.

Also available via the World Wide Web.

Martin:2005:PET

[MS05] David Martin and Andrei Serjantov, editors. Privacy Enhancing Technologies: 4th International Workshop, PET 2004, Toronto, Canada, May 26–28, 2004. Revised Selected Papers, volume 3424 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26203-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 P49 2004.

URL http://www.springerlink.
com/content/978-3-540-
26203-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3424.

Margaria:2006:LAF

[MS06a] Tiziana Margaria and Bernhard Steffen, editors. Leveraging Applications of Formal Methods: First International Symposium, ISoLA 2004, Paphos, Cyprus, October 30 — November2, 2004, Revised Selected Papers, volume 4313 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48928-2 (print), 3-540-48929-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48929-0.

Missaoui:2006:FCA

[MS06b] Rokia Missaoui and Jürg Schmidt, editors. Formal Concept Analysis: 4th International Conference, ICFCA 2006, Dresden, Germany, February 13–17, 2006. Proceedings, volume 3874 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32203-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32203-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3874.

Mittermeir:2008:IES

[MS08a] Roland T. Mittermeir and Maciej M. Sysło, editors. Informatics Education — Supporting Computational Thinking: Third International Conference on Informatics in Secondary Schools — Evolution and Perspectives, ISSEP 2008 Torun Poland, July 1–4, 2008 Proceedings, volume 5090 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69923-6 (print), 3-540-69924-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69924-8.

Monien:2008:AGT

[MS08b] Burkhard Monien and Ulf-Peter Schroeder, editors. Algorithmic Game Theory: First International Symposium, SAGT 2008, Paderborn, Germany, April 30–May 2, 2008. Proceedings, volume 4997 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79308-9 (print), 3-540-79309-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79309-0.

Mangard:2010:CHE

[MS10a] Stefan Mangard and François-Xavier Standaert, editors. Cryptographic Hardware and Embedded Systems, CHES 2010: 12th International Workshop, Santa Barbara, USA, August 17– 20, 2010. Proceedings, volume 6225 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15030-6 (print), 3-642-15031-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15031-9.

Margaria:2010:LAFa

[MS10b] Tiziana Margaria and Bernhard Steffen, editors. Leveraging Applications of Formal Methods, Verification, and Validation: 4th International Symposium on Leveraging Applications, ISoLA 2010, Heraklion, Crete, Greece, October 18–21, 2010, Proceedings, Part I, volume 6415 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16557-5 (print), 3-642-16558-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16558-0.

Margaria:2010:LAFb

[MS10c] Tiziana Margaria and Bernhard Steffen, editors. Leveraging Applications of Formal Methods, Verification, and Validation: 4th International Symposium on Leveraging Applications, ISoLA 2010, Heraklion, Crete, Greece, October 18–21, 2010, Proceedings, Part II, volume 6416 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16560-5 (print), 3-642-16561-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16561-0.

May:2010:UKD

[MS10d] Michael May and Lorenza Saitta, editors. Ubiquitous Knowledge Discovery: Challenges, Techniques, Applications, volume 6202 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16391-2 (print), 3-642-16392-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16392-0.

Moulton:2010:ABI

[MS10e] Vincent Moulton and Mona Singh, editors. Algorithms in Bioinformatics: 10th International Workshop, WABI 2010, Liverpool, UK, September 6–8, 2010. Proceedings, volume 6293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15293-7 (print), 3-642-15294-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15294-8.

Morazan:2011:IAF

[MS11a] Marco T. Morazán and Sven-Bodo Scholz, editors. Implementation and Application of Functional Languages: 21st International Symposium, IFL 2009, South Orange, NJ, USA, September 23–25, 2009, Revised Selected Papers, volume 6041 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16477-3 (print), 3-642-16478-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16478-1.

Murlak:2011:MFC

[MS11b] Filip Murlak and Piotr Sankowski, editors. Mathematical Foundations of Computer Science 2011: 36th International Symposium, MFCS 2011, Warsaw, Poland, August 22–26, 2011. Proceedings, volume 6907 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22992-1 (print), 3-642-22993-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        22993-0.

Madhusudan:2012:CAV

[MS12a] P. Madhusudan and Sanjit A. Seshia, editors. Computer Aided Verification: 24th International Conference, CAV 2012, Berkeley, CA, USA, July 7–13, 2012 Proceedings, volume 7358 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31423-6 (print), 3-642-31424-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31424-7.

Margaria:2012:LAF

[MS12b] Tiziana Margaria and Bernhard Steffen, editors. Leveraging Applications of Formal Methods, Verification and Validation. Applications and Case Studies: 5th International Symposium, ISoLA 2012, Heraklion, Crete, Greece, October 15–18, 2012, Proceedings, Part II, volume 7610 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34031-8 (print), 3-642-34032-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34032-1.

Mine:2012:SAI

[MS12c] Antoine Miné and David Schmidt, editors. Static Analysis: 19th International Symposium, SAS 2012, Deauville, France, September 11–13, 2012. Proceedings, volume 7460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33124-6 (print), 3-642-33125-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33125-1.

Mira:1999:IWCa

[MSA99a] José Mira and Juan V.

Sánchez-Andrés, editors. International Work Conference on Artificial and Natural Neural Networks, Alicante, Spain, June 2–4, 1999: Proceedings. Part 1. Foundations and Tools for Neural Modeling, volume 1606 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66069-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP363.3 .I586 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1606.htm;
 http://www.springerlink.
com/content/978-3-540-
66069-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1606.

Mira:1999:IWCb

[MSA99b] José Mira and Juan V. Sánchez-Andrés, editors. International Work Conference on Artificial and Natural Neural Networks, Alicante, Spain, June 2–4, 1999: Proceedings. Part 2. Engineering Applications of Bio-Inspired Artificial Neural Networks, volume 1607 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66068-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP363.3 .I586 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1607.htm;
 http://www.springerlink.

com/content/978-3-540-
66068-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1607.

Maziero:2005:DCS

[MSAdAS05] Carlos Alberto Maziero, João Gabriel Silva, Aline Maria Santos Andrade, and Flávio Morais de Assis Silva, editors. Dependable Computing: Second Latin-American Symposium, LADC 2005, Salvador, Brazil, October 25–28, 2005. Proceedings, volume 3747 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29572-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29572-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3747.

Montanari:2007:TGC

[MSB07] Ugo Montanari, Donald Sannella, and Roberto Bruni, editors. Trustworthy Global Computing: Second Symposium, TGC 2006, Lucca, Italy, November 7–9, 2006, Revised Selected Papers, volume 4661 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75333-8 (print), 3-540-75336-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75336-0.

Magnusson:2012:HAI

[MSB12] Charlotte Magnusson, Delphine Szymczak, and Stephen Brewster, editors. Haptic and Audio Interaction Design: 7th International Conference, HAID 2012, Lund, Sweden, August 23–24, 2012. Proceedings, volume 7468 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32795-8 (print), 3-642-32796-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32796-4.

Mizoguchi:2006:SWA

[MSG06] Riichiro Mizoguchi, Zhongzhi Shi, and Fausto Giunchiglia, editors. The Semantic Web — ASWC 2006:

First Asian Semantic Web Conference, Beijing, China, September 3–7, 2006. Proceedings, volume 4185 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38329-8 (print), 3-540-38331-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38331-4.

Maicher:2007:LST

[MSG07] Lutz Maicher, Alexander Sigel, and Lars Marius Garshol, editors. Leveraging the Semantics of Topic Maps: Second International Conference on Topic Maps Research and Applications, TMRA 2006, Leipzig, Germany, October 11–12, 2006, Revised Selected Papers, volume 4438 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71944-X (print), 3-540-71945-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        71945-8.

Mariani:1985:LAN

[MSH85] J. A. Mariani, W. D. Shepherd, and David Hutchison, editors. Local area networks: an advanced course, Glasgow, July 11–22, 1983 proceedings, volume 184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15191-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.7 .L64 1985. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0184.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        15191-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=184.

Motoda:2005:DSI

[MSH05] Hiroshi Motoda, Tobias Scheffer, and Achim Hoffmann, editors. Discovery science: 8th international conference, DS 2005, Singapore, October 8–11, 2005: proceedings, volume 3735 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29230-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q174 .D57 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29230-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3735.

Miculan:2008:TPP

[MSH08] Marino Miculan, Ivan Scagnetto, and Furio Honsell, editors. Types for Proofs and Programs: International Conference, TYPES 2007, Cividale des Friuli, Italy, May 2–5, 2007 Revised Selected Papers, volume 4941 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68084-5 (print), 3-540-68103-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68103-8.

Moraglio:2012:GPE

[MSK^{+}12] Alberto Moraglio, Sara Silva, Krzysztof Krawiec, Penousal Machado, and Carlos Cotta, editors.

Genetic Programming: 15th European Conference, EuroGP 2012, Málaga, Spain, April 11–13, 2012. Proceedings, volume 7244 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29138-4 (print), 3-642-29139-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29139-5.

Marik:2002:MAS

[MŠKL02] Vladimir Marik, Olga Štěpánková, Hana Krautwurmová, and Michael Luck, editors. Multi-agent systems and applications II: 9th ECCAI-ACAI/EASSS 2001, AE-MAS 2001, HoloMAS 2001: Selected Revised Papers, volume 2322 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43377-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M86 2002. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t2322.htm;
http://www.springerlink.
com/content/978-3-540-
43377-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2322.

Muller-Schloer:2010:ACS

[MSKY10] Christian Müller-Schloer, Wolfgang Karl, and Sami Yehia, editors. Architecture of Computing Systems — ARCS 2010: 23rd International Conference, Hannover, Germany, February 22– 25, 2010. Proceedings, volume 5974 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11949-2 (print), 3-642-11950-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11950-7.

Mamoulis:2009:AST

[MSP^{+}09] Nikos Mamoulis, Thomas Seidl, Torben Bach Pedersen, Kristian Torp, and Ira Assent, editors. Advances in Spatial and Temporal Databases: 11th International Symposium, SSTD 2009 Aalborg, Denmark, July 8– 10, 2009 Proceedings, volume 5644 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02981-7 (print), 3-642-02982-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02982-0.

Muller:1999:IAV

[MSR99] J. P. Müller, Munindar P. Singh, and Anand S. Rao, editors. Intelligent agents V: agent theories, architectures, and languages: 5th International Workshop, ATAL’98: Paris, France, July 4–7, 1998: Proceedings, volume 1555 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65713-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1555. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1555.htm;
 http://www.springerlink.
com/content/978-3-540-

65713-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1555.

Margaria:1998:SVT

[MSRP98] Tiziana Margaria, Bernhard Steffen, Roland Rückert, and Joachim Posegga, editors. Services and visualization: towards user-friendly design: ACoS’98, VI-SUAL’98, AIN’97 Selected Papers, volume 1385 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64367-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.U83A29 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1385.
htm;   http://www.springerlink.
com/content/978-3-540-
64367-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1385.

Mogensen:2002:ECC

[MSS02] Torben A. Mogensen, David A. Schmidt, and Hal Sudborough, editors. The Essence of Computation. Complexity, Analysis, Transformation: Essays Dedicated to Neil D. Jones, volume 2566 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00326-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 471 pp. LCCN QA76.6 .E86 2002. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2566.
htm;   http://www.springerlink.
com/content/978-3-540-
00326-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2566.

Also available via the World Wide Web.

Menasalvas:2003:AWI

[MSS03a] Ernestina Menasalvas, Javier Segovia, and Piotr S. Szczepaniak, editors. Advances in Web intelligence: First International Atlantic Web Intelligence Conference, AWIC 2003, Madrid, Spain, May 5–6, 2003: Proceedings, volume 2663 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40124-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN TK5105.888 .I5726 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2663.htm;
 http://www.springerlink.
com/content/978-3-540-
40124-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2663.

Also available via the World Wide Web.

Meyer:2003:AMH

[MSS03b] Ulrich Meyer, Peter Sanders, and Jop Sibeyn, editors. Algorithms for Memory Hierarchies: Advanced Lectures, volume 2625 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00883-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xvii + 428 pp. LCCN QA76.9.A43 A435 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2625.
htm;   http://www.springerlink.
com/content/978-3-540-
00883-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2625.

Also available via the World Wide Web.

Motta:2004:EKA

[MSS04] Enrico Motta, Nigel R. Shadbolt, and Arthur Stutt, editors. Engineering Knowledge in the Age of the Semantic Web: 14th International Conference, EKAW 2004, Whittlebury Hall, UK, October 5–8, 2004: Proceedings, volume 3257 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23340-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3257.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23340-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3257;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b101291.

Moro:2005:APPa

[MSS05a] Gianluca Moro, Claudio Sartori, and Munindar P. Singh, editors. Agents and peer-to-peer computing: second international workshop, AP2PC 2003, Melbourne, Australia, July 14, 2003, revised and invited papers, volume 2872 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24053-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
24053-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2872.

Murray-Smith:2005:SLF

[MSS05b] Roderick Murray-Smith and Robert Shorten, editors. Switching and learning in feedback systems: European summer school on multiagent control, Maynooth, Ireland, September 8– 10, 2003. Revised lectures and selected papers, volume 3355 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24457-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ216 .E87 2003.

URL http://www.springerlink.
com/content/978-3-540-
24457-8;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3355;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105497.

Marques-Silva:2007:TAS

[MSS07] João Marques-Silva and Karem A. Sakallah, editors. Theory and Applications of Satisfiability Testing – SAT 2007: 10th International Conference, Lisbon, Portugal, May 28–31, 2007. Proceedings, volume 4501 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72787-6 (print), 3-540-72788-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72788-0.

Mu:2008:ISP

[MSS08] Yi Mu, Willy Susilo, and Jennifer Seberry, editors. Information Security and Privacy: 13th Australasian Conference, ACISP 2008, Wollongong, Australia, July 7–9, 2008. Proceedings, volume 5107 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69971-6 (print), 3-540-70500-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70500-0.

Marchetti-Spaccamela:2011:TPA

[MSS11] Alberto Marchetti-Spaccamela and Michael Segal, editors. Theory and Practice of Algorithms in (Computer) Systems: First International ICST Conference, TAPAS 2011, Rome, Italy, April 18– 20, 2011. Proceedings, volume 6595 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19753-1 (print), 3-642-19754-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19754-3.

Margaria-Steffen:2012:LAF

[MSS12] Tiziana Margaria-Steffen and Bernhard Steffen, editors. Leveraging applications of formal methods, verification and validation. Technologies for mastering change: 5th International Symposium, ISoLA 2012, Heraklion, Crete, Greece, October 15–18, 2012, Proceedings. Part I, volume 7609 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34025-3, 3-642-34026-1 (e-book). LCCN QA76.9.F67 I86 2012.

URL http://www.springerlink.
com/content/978-3-642-
34025-3;   http://www.
springerlink.com/content/
978-3-642-34026-0.

Marik:1992:ATA

[MST92] V. Marik, O. Stepankova, and Robert Trappl, editors. Advanced topics in artificial intelligence: international summer school, Prague, Czechoslovakia, July 6– 17, 1992: proceedings, volume 617 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55681-8 (Berlin), 0-387-55681-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A343 1992.

URL http://link.springer-
ny.com/link/service/

series/0558/tocs/t0617.
htm;   http://www.springerlink.
com/content/978-3-540-
55681-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=617.

Mayr:1995:GTC

[MST95] Ernst Mayr, Gunther Schmidt, and Gottfried Tinhofer, editors. Graphtheoretic concepts in computer science: 20th International Workshop, WG ’94, Herrsching, Germany, June 16–18, 1994: proceedings, volume 903 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59071-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G715 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0903.htm;
http://www.springerlink.
com/content/978-3-540-
59071-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=903.

Mana:2012:ANN

[MST12] Nadia Mana, Friedhelm Schwenker, and Edmondo Trentin, editors. Artificial Neural Networks in Pattern Recognition: 5th INNS IAPR TC 3 GIRPR Workshop, ANNPR 2012, Trento, Italy, September 17–19, 2012. Proceedings, volume 7477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33211-0 (print), 3-642-33212-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33212-8.

Muller-Schloer:2004:OPC

[MSUB04] Christian Müller-Schloer, Theo Ungerer, and Bernhard Bauer, editors. Organic and Pervasive Computing—ARCS 2004: International Conference on Architecture of Computing Systems, Augsburg, Germany, March 23–26, 2004: Proceedings, volume 2981 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21238-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A73 A736 2004.

URL http://link.springer-
ny.com/link/service/

series/0558/tocs/t2981.
htm;   http://www.springerlink.
com/content/978-3-540-
21238-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2981;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95942.

Meng:2002:AWA

[MSW02] Xiaofeng Meng, Jianwen Su, and Yujun Wang, editors. Advances in Web-Age Information Management: Third International Conference, WAIM 2002, Beijing, China, August 11–13, 2002. Proceedings, volume 2419 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44045-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 W325 2002 Bar. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2419.htm;
http://www.springerlink.
com/content/978-3-540-
44045-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2419.

Maybury:2005:ITI

[MSW05] Mark T. Maybury, Oliviero Stock, and Wolfgang Wahlster, editors. Intelligent technologies for interactive entertainment: first international conference, INTETAIN 2005, Madonna di Campiglio, Italy, November 30– December 2, 2005: proceedings, volume 3814 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30509-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.E57 I65 2005.

URL http://www.springerlink.
com/content/978-3-540-
30509-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3814.

Martin:2007:MSX

[MSW07] Ralph Martin, Malcolm Sabin, and Joab Winkler, editors. Mathematics of Surfaces XII: 12th IMA International Conference, Sheffield, UK, September 4–6, 2007. Proceedings, volume 4647 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73842-8 (print), 3-540-73843-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73843-5.

Matsumoto:2006:CPO

[MSWZ06] Yuji Matsumoto, Richard W.

Sproat, Kam-Fai Wong, and Min Zhang, editors. Computer Processing of Oriental Languages. Beyond the Orient: The Research Challenges Ahead: 21st International Conference, IC-CPOL 2006, Singapore, December 17–19, 2006. Proceedings, volume 4285 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49667-X (print), 3-540-49668-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49668-7.

Mourrain:2010:AGM

[MSX10] Bernard Mourrain, Scott Schaefer, and Guoliang Xu, editors. Advances in Geometric Modeling and Processing: 6th International Conference, GMP 2010, Castro Urdiales, Spain, June 16–18, 2010. Proceedings, volume 6130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13410-6 (print), 3-642-13411-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13411-1.

Moukas:2000:AME

[MSY00] Alexandros Moukas, Carles Sierra, and Fredrik Ygge, editors. Agent mediated electronic commerce II: towards nextgeneration agent-based electronic commerce systems: Proceedings of a workshop held during the Sixteenth International Joint Conference on Artificial Intelligence in Stockholm, Sweden in July 1999, volume 1788 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67773-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.32 .A34 2000.

URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t1788.
htm;   http://www.springerlink.
com/content/978-3-540-
67773-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1788.

Marinaro:2008:DBN

[MSY08] Maria Marinaro, Silvia Scarpetta, and Yoko Yamaguchi, editors. Dynamic Brain — from Neural Spikes to Behaviors: 12th International Summer School on Neural Networks, Erice, Italy, December 5–12, 2007, Revised Lectures, volume 5286 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88852-7 (print), 3-540-88853-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88853-6.

Marik:1990:AIH

[MSZ90] V. Marik, O. Stepankova, and Z. Zdrahal, editors. Artificial intelligence in higher education: CEPES-UNESCO International Symposium, Prague, CSFR, October 23–25, 1989: proceedings, volume 451 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52952-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.43 .C46 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0451.htm;
 http://www.springerlink.
com/content/978-0-387-
52952-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=451.

Mandoiu:2008:BRA

[MSZ08] Ion Măndoiu, Raj Sunderraman, and Alexander Zelikovsky, editors. Bioinformatics Research and Applications: Fourth International Symposium, ISBRA 2008, Atlanta, GA, USA, May 6–9, 2008. Proceedings, volume 4983 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79449-2 (print), 3-540-79450-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79450-9.

Marik:2009:HMA

[MSZ09] Vladimír Mařík, Thomas Strasser, and Alois Zoitl, editors. Holonic and Multi-Agent Systems for Manufacturing: 4th International Conference on Industrial Applications of Holonic and Multi-Agent Systems, HoloMAS 2009, Linz, Austria, August 31 — September 2, 2009. Proceedings, volume 5696 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03666-X (print), 3-642-03668-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03668-2.

Miranker:1986:ASC

[MT86] Willard L. Miranker and Richard A. Toupin, editors. Accurate scientific computations: symposium, Bad Neuenahr, FRG, March 12–14, 1985: proceedings, volume 235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CODEN LNCSD9. ISBN 0-387-16798-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .A231 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0235.htm;
 http://www.springerlink.
com/content/978-0-387-
16798-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=235.

Symposium sponsored by IBM Deutschland.

Meyer:1989:LBS

[MT89] Albert R. Meyer and M. A. Taitslin, editors. Logic at Botik ’89: symposium on logical foundations of computer science, Pereslavl-Zalessky, USSR, July 3–8, 1989: proceedings, volume 363 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51237-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .B67 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0363.htm;
 http://www.springerlink.
com/content/978-0-387-
51237-2;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=363.

Mavronicolas:1997:DAI

[MT97] Mario Mavronicolas and Philippas Tsigas, editors. Distributed algorithms: 11th International Workshop, WDAG ’97, Saarbrucken, Germany, September 24– 26, 1997: proceedings, volume 1320 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63575-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I597 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1320.htm;
http://www.springerlink.
com/content/978-3-540-
63575-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1320.

Meinel:1999:SAS

[MT99a] Christoph Meinel and Sophie Tison, editors. STACS 99: 16th annual Symposium on Theoretical Aspects of Computer Science, Trier, Germany, March 4–6, 1999: Proceedings, volume 1563 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65691-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S958 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1563.htm;
 http://www.springerlink.
com/content/978-3-540-
65691-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1563.

Mohania:1999:DWK

[MT99b] Mukesh Mohania and A. Min Tjoa, editors. Data warehousing and knowledge discovery: First International Conference, DaWaK’99, Florence, Italy, August 30– September 1, 1999: Proceedings, volume 1676 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66458-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1676.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1676.
htm;   http://www.springerlink.
com/content/978-3-540-
66458-1;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1676.

Moller:1999:PHF

[MT99c] Bernhard Möller and J. V. (John V.) Tucker, editors. Prospects for hardware foundations: ESPRIT Working Group 8533: NADA — new hardware design methods, survey chapters, volume 1546 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65461-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7888.3.P766 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1546.htm;
http://www.springerlink.
com/content/978-3-540-
65461-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1546.

Marinaro:2002:NNI

[MT02a] M. Marinaro and Roberto Tagliaferri, editors. Neural Nets: 13th Italian Workshop on Neural Nets, WIRN VI-ETRI 2002, Vietri sul Mare, Italy, May 30– June 1, 2002. Revised Papers, volume 2486 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44265-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I58917 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2486.htm;
 http://www.springerlink.
com/content/978-3-540-
44265-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2486.

Meyer:2002:IAV

[MT02b] John-Jules Ch Meyer and Milind Tambe, editors. Intelligent agents VIII: agent theories, architectures, and languages: 8th International Workshop, ATAL 2001, Seattle, WA, USA, August 1–3, 2001: Revised Papers, volume 2333 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43858-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 A83 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2333.

htm;   http://www.springerlink.
com/content/978-3-540-
43858-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2333.

Meersman:2003:MMIb

[MT^{+}03] Robert Meersman, Zahir Tari, et al., editors. On The Move to Meaningful Internet Systems 2003: OTM 2003 Workshops: OTM confederated International Workshops, HCI-SWWA, IPW, JTRES, WORM, WMS, and WRSM 2003, Catania, Sicily, Italy, November 3–7, 2003: Proceedings, volume 2889 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20494-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 O488 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2889.
htm;   http://www.springerlink.
com/content/978-3-540-
20494-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2889;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94345.

Marcinkowski:2004:CSL

[MT04a] Jerzy Marcinkowski and Andrzej Tarlecki, editors. Computer Science Logic: 18th International Workshop, CSL 2004, 13th Annual Conference of the EACSL, Karpacz, Poland, September 20– 24, 2004: Proceedings, volume 3210 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23024-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 S23 2004.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3210.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        23024-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3210;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100120.

McDonald:2004:AIR

[MT04b] Sharon McDonald and John Tait, editors. Advances in Information Retrieval: 26th European Conference on IR Research, ECIR 2004, Sunderland, UK, April 5–7, 2004: Proceedings, volume 2997 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21382-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 E92 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2997.
htm;   http://www.springerlink.
com/content/978-3-540-
21382-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2997;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96895.

Meersman:2004:MMIa

[MT04c] Robert Meersman and Zahir Tari, editors. On the Move to Meaningful Internet Systems 2004: CoopIS, DOA, and ODBASE, volume 3290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23663-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3290.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        23663-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3290;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b102173.

Meersman:2004:MMIb

[MT04d] Robert Meersman and Zahir Tari, editors. On the Move to Meaningful Internet Systems 2004: CoopIS, DOA, and ODBASE, volume 3291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23662-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3291.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23662-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3291;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b102176.

Meersman:2005:MMIa

[MT05a] Robert Meersman and Zahir Tari, editors. On the move to meaningful Internet systems 2005: CoopIS, DOA, and ODBASE: OTM Confederated International Conferences CoopIS, DOA, and ODBASE 2005, Agia Napa, Cyprus, October 31–November 4, 2005: proceedings, volume 3760–3761 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29736-7, 3-540-29738-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 O48 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29736-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3760-
-3761.

Morari:2005:HSC

[MT05b] Manfred Morari and Lothar Thiele, editors. Hybrid systems: computation and control: 8th international workshop, HSCC 2005, Zürich, Switzerland, March 9– 11, 2005: Proceedings, volume 3414 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25108-1 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38 .H73 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25108-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3414;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106766.

Meersman:2006:MMIa

[MT06a] Robert Meersman and Zahir Tari, editors. On the Move to Meaningful Internet Systems 2006: CoopIS, DOA, GADA, and ODBASE: OTM Confederated International Conferences, CoopIS, DOA, GADA, and ODBASE 2006, Montpellier, France, October 29 — November 3, 2006. Proceedings, Part I, volume 4275 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48287-3 (print), 3-540-48289-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        48289-5.

Meersman:2006:MMIb

[MT06b] Robert Meersman and Zahir Tari, editors. On the Move to Meaningful Internet Systems 2006: CoopIS, DOA, GADA, and ODBASE: OTM Confederated International Conferences, CoopIS, DOA, GADA, and ODBASE 2006, Montpellier, France, October 29 — November 3, 2006. Proceedings, Part II, volume 4276 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48274-1 (print), 3-540-48283-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48283-3.

Masuzawa:2007:SSS

[MT07a] Toshimitsu Masuzawa and Sébastien Tixeuil, editors. Stabilization, Safety, and Security of Distributed Systems: 9th International Symposium, SSS 2007 Paris, France, November 14– 16, 2007 Proceedings, volume 4838 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76626-X (print), 3-540-76627-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76627-8.

Meersman:2007:MMIa

[MT07b] Robert Meersman and Zahir Tari, editors. On the Move to Meaningful Internet Systems 2007: CoopIS, DOA, ODBASE, GADA, and IS: OTM Confederated International Conferences CoopIS, DOA, ODBASE, GADA, and IS 2007, Vilamoura, Portugal, November 25– 30, 2007, Proceedings, Part I, volume 4803 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76846-7 (print), 3-540-76848-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76848-7.

Meersman:2007:MMIb

[MT07c] Robert Meersman and Zahir Tari, editors. On the Move to Meaningful Internet Systems 2007: CoopIS, DOA, ODBASE, GADA, and IS: OTM Confederated International Conferences CoopIS, DOA, ODBASE, GADA, and IS 2007, Vilamoura, Portugal, November 25– 30, 2007, Proceedings, Part II, volume 4804 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76835-1 (print), 3-540-76843-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76843-2.

Meersman:2008:MMIa

[MT08a] Robert Meersman and Zahir Tari, editors. On the Move to Meaningful Internet Systems: OTM 2008: OTM 2008 Confederated International Conferences, CoopIS, DOA, GADA, IS, and ODBASE 2008, Monterrey, Mexico, November 9–14, 2008, Proceedings, Part I, volume 5331 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88870-5 (print), 3-540-88871-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88871-0.

Meersman:2008:MMIb

[MT08b] Robert Meersman and Zahir Tari, editors. On the Move to Meaningful Internet Systems: OTM 2008: OTM 2008 Confederated International Conferences, CoopIS, DOA, GADA, IS, and ODBASE 2008, Monterrey, Mexico, November 9–14, 2008, Proceedings, Part II, volume 5332 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88872-1 (print), 3-540-88873-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88873-4.

Meier:2008:DAI

[MT08c] René Meier and Sotirios Terzis, editors. Distributed Applications and Interoperable Systems: 8th IFIP WG 6.1 International Conference, DAIS 2008, Oslo, Norway, June 4–6, 2008.

Proceedings, volume 5053 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68639-8 (print), 3-540-68642-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68642-2.

Macintosh:2009:EPF

[MT09a] Ann Macintosh and Efthimios Tambouris, editors. Electronic Participation: First International Conference, ePart 2009 Linz, Austria, September 1–3, 2009 Proceedings, volume 5694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03780-1 (print), 3-642-03781-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03781-8.

Magnenat-Thalmann:2009:MPH

[MT09b] Nadia Magnenat-Thalmann, editor. Modelling the Physiological Human: 3D Physiological Human Workshop, 3DPH 2009, Zermatt, Switzerland, November 29 — December 2, 2009. Proceedings, volume 5903 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10468-1 (print), 3-642-10470-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10470-1.

Majumdar:2009:HSC

[MT09c] Rupak Majumdar and Paulo Tabuada, editors. Hybrid Systems: Computation and Control: 12th International Conference, HSCC 2009, San Francisco, CA, USA, April 13–15, 2009. Proceedings, volume 5469 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00601-9 (print), 3-642-00602-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00602-9.

Meersman:2004:MMIc

[MTC04] Robert Meersman, Zahir Tari, and Angelo Corsaro, editors. On the Move to Meaningful Internet Systems 2004: OTM 2004 Workshops, volume 3292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23664-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3292.htm;
 http://www.springerlink.
com/content/978-3-540-
23664-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3292;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b102133.

Monmarche:2008:AEI

[MTC^{+}08] Nicolas Monmarché, El-Ghazali Talbi, Pierre Collet, Marc Schoenauer, and Evelyne Lutton, editors. Artificial Evolution: 8th International Conference, Evolution Artificielle, EA 2007, Tours, France, October 29–31, 2007, Revised Selected Papers, volume 4926 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79304-6 (print), 3-540-79305-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79305-2.

Martinez-Trinidad:2011:PRT

[MTCOBH11] José Francisco Martínez-Trinidad, Jesús Ariel Carrasco-Ochoa, Cherif Ben-Youssef Brants, and Edwin Robert Hancock, editors. Pattern Recognition: Third Mexican Conference, MCPR 2011, Cancun, Mexico, June 29 — July 2, 2011. Proceedings, volume 6718 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21586-6 (print), 3-642-21587-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21587-2.

Martinez-Trinidad:2010:APR

[MTCOK10] José Francisco Martínez-Trinidad, Jesús Ariel Carrasco-Ochoa, and Josef Kittler, editors. Advances in Pattern Recognition: Second Mexican Conference on Pattern Recognition, MCPR 2010, Puebla, Mexico, September 27–29, 2010. Proceedings, volume 6256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15991-5 (print), 3-642-15992-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15992-3.

Meersman:2005:MMIb

[MTH05] Robert Meersman, Zahir Tari, and Pilar Herrero, editors. On the move to meaningful Internet systems 2005: OTM 2005 Workshops: OTM Confederated International Workshops and Posters: AWe-SOMe, CAMS, GADA, MIOS+INTEROP, ORM, PhDS, SeBGIS, SWWS, and WOSE 2005, Agia Napa, Cyprus, October 31–November 4, 2005: proceedings, volume 3762 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29739-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 O483 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        29739-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3762.

Meersman:2006:MMIc

[MTH06a] Robert Meersman, Zahir Tari, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems 2006: OTM 2006 Workshops: OTM Confederated International Workshops and Posters, AWeSOMe, CAMS, COMINF, IS, KSinBIT, MIOS-CIAO, MONET, OnToContent, ORM, PerSys, OTM Academy Doctoral Consortium, RDDS, SWWS, and SeBGIS 2006, Montpellier, France, October 29 — November 3, 2006. Proceedings, Part I, volume 4277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48269-5 (print), 3-540-48272-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        48272-7.

Meersman:2006:MMId

[MTH06b] Robert Meersman, Zahir Tari, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems 2006: OTM 2006 Workshops: OTM Confederated International Workshops and Posters, AWeSOMe, CAMS, COMINF, IS, KSinBIT, MIOS-CIAO, MONET, OnToContent, ORM, PerSys, OTM Academy Doctoral Consortium, RDDS, SWWS, and SeBGIS 2006, Montpellier, France, October 29 — November 3, 2006. Proceedings, Part II, volume 4278 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48273-3 (print), 3-540-48276-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48276-5.

Meersman:2007:MMIc

[MTH07a] Robert Meersman, Zahir Tari, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems 2007: OTM 2007 Workshops: OTM Confederated International Workshops and Posters, AWeSOMe, CAMS, OTM Academy Doctoral Consortium, MONET, OnToContent, ORM, PerSys, PPN, RDDS, SSWS, and SWWS 2007, Vilamoura, Portugal, November 25–30, 2007, Proceedings, Part I, volume 4805 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76887-4 (print), 3-540-76888-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76888-3.

Meersman:2007:MMId

[MTH07b] Robert Meersman, Zahir Tari, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems 2007: OTM 2007 Workshops: OTM Confederated International Workshops and Posters, AWeSOMe, CAMS, OTM Academy Doctoral Consortium, MONET, OnToContent, ORM, PerSys, PPN, RDDS, SSWS, and SWWS 2007, Vilamoura, Portugal, November 25–30, 2007, Proceedings, Part II, volume 4806 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76889-0 (print), 3-540-76890-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76890-6.

Meersman:2008:MMIc

[MTH08] Robert Meersman, Zahir Tari, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems: OTM 2008 Workshops: OTM Confederated International Workshops and Posters, ADI, AWeSoMe, COMBEK, EI2N, IWSSA, MONET, OnToContent + QSI, ORM, PerSys, RDDS, SEMELS, and SWWS 2008, Monterrey, Mexico, November 9–14, 2008. Proceedings, volume 5333 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88874-8 (print), 3-540-88875-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        88875-8.

Muggleton:2012:ILP

[MTNL12] Stephen H. Muggleton, Alireza Tamaddoni-Nezhad, and Francesca A. Lisi, editors. Inductive Logic Programming: 21st International Conference, ILP 2011, Windsor Great Park, UK, July 31 — August 3, 2011, Revised Selected Papers, volume 7207 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31950-5 (print), 3-642-31951-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31951-8.

Matsuoka:1999:COO

[MTO99] Satoshi Matsuoka, Marydell Tholburn, and Rodney R. Oldehoeft, editors. Computing in object-oriented parallel environments: Third International Symposium, ISCOPE 99, San Francisco, CA, USA, December 8–10, 1999: Proceedings, volume 1732 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66818-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1732. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1732.htm;
 http://www.springerlink.
com/content/978-3-540-
66818-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1732.

Martinez-Trinidad:2006:PPR

[MTOK06] José Francisco Martínez-Trinidad, Jesús Ariel Carrasco Ochoa, and Josef Kittler, editors. Progress in Pattern Recognition, Image Analysis and Applications: 11th Iberoamerican Congress in Pattern Recognition, CIARP 2006 Cancun, Mexico, November 14–17, 2006 Proceedings, volume 4225 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46556-1 (print), 3-540-46557-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46557-7.

Meersman:2002:MMI

[MTP^{+}02] Robert Meersman, Zahir Tari, Michael Papazoglou, Stefano Spaccapietra, Calton Pu, Rachid Guerraoui, Joe Loyall, Douglas Schmidt, Karl Aberer, and Ling Liu, editors. On the Move to Meaningful Internet Systems 2002: CoopIS, DOA, and ODBASE: Confederated International Conferences: CoopIS, DOA, and ODBASE 2002. Proceedings, volume 2519 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00106-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .I54 2002; QA76.9.D3 I5814 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2519.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00106-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2519.

Meersman:2003:MMIa

[MTS^{+}03] Robert Meersman, Zahir Tari, Douglas C. Schmidt, et al., editors. On The Move to Meaningful Internet Systems 2003: CoopIS, DOA, and ODBASE: OTM confederated International Conferences CoopIS, DOA, and ODBASE 2003, Catania, Sicily, Italy, November 3–7, 2003: Proceedings, volume 2888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20498-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2888.htm;
http://www.springerlink.
com/content/978-3-540-
20498-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2888;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94348.

Magnenat-Thalmann:1998:MMC

[MTT98] Nadia Magnenat-Thalmann and Daniel Thalmann, editors. Modelling and Motion Capture Techniques for Virtual Environments: International Workshop, CAPTECH’98, Geneva, Switzerland, November 126-27, 998. Proceedings, volume 1537 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65353-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QP301.C25 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1537.
htm;   http://www.springerlink.
com/content/978-3-540-
65353-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1537.

Moon:2009:PAN

[MTU09] Sue B. Moon, Renata Teixeira, and Steve Uhlig, editors. Passive and Active Network Measurement: 10th International Conference, PAM 2009, Seoul, Korea, April 1–3, 2009. Proceedings, volume 5448 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00974-3 (print), 3-642-00975-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00975-4.

Masulli:2009:CIM

[MTV09] Francesco Masulli, Roberto Tagliaferri, and Gennady M. Verkhivker, editors. Computational Intelligence Methods for Bioinformatics and Biostatistics: 5th International Meeting, CIBB 2008 Vietri sul Mare, Italy, October 3–4, 2008 Revised Selected Papers, volume 5488 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02503-X (print), 3-642-02504-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02504-4.

Molva:2005:SPA

[MTW05] Refik Molva, Gene (Gene)

Tsudik, and Dirk Westhoff, editors. Security and privacy in adhoc and sensor networks: second European Workshop, ESAS 2005, Visegrad, Hungary, July 13– 14, 2005: revised selected papers, volume 3813 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30912-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5102.85 .E97 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        30912-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3813.

Mohr:2006:RAP

[MTWD06] Bernd Mohr, Jesper Larsson Träff, Joachim Worringen, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 13th European PVM/ MPI User’s Group Meeting Bonn, Germany, September 17–20, 2006 Proceedings, volume 4192 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39110-X (print), 3-540-39112-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        39112-8.

Mueller:2001:HLP

[Mue01] Frank Mueller, editor. High-level parallel programming models and supportive environments: 6th International Workshop, HIPS 2001, San Francisco, CA, USA, April 23, 2001: Proceedings, volume 2026 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41944-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .I586 2001; QA267.A1.L43 no.2026. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2026.htm;
 http://www.springerlink.
com/content/978-3-540-
41944-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2026.

Mueller:2006:ETI

[Mue06] Guenter Mueller, editor.

Emerging Trends in Information and Communication Security: International Conference, ETRICS 2006, Freiburg, Germany, June 6–9, 2006. Proceedings, volume 3995 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34640-6 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34640-1.

Muggleton:1997:ILP

[Mug97] Stephen Muggleton, editor. Inductive logic programming: 6th International Workshop, ILP-96, Stockholm, Sweden, August 1996: Selected Papers, volume 1314 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63494-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63.I52 1997.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1314.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63494-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1314.

Muhlbacher:1975:GFJ

[Müh75] Jörg Mühlbacher, editor. GI — 5. [i.e., funfte] Jahrestagung: Dortmund, 8.–10. Oktober 1975. (German) [GI: Fifth Annual Conference, Dortmund, October 8– 10, 1975], volume 34 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CO-DEN LNCSD9. ISBN 0-387-07410-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.34.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0034.
htm;   http://www.springerlink.
com/content/978-0-387-
07410-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=34.

English or German.

Mulkers:1993:LDS

[Mul93]  Anne Mulkers.  Live

data structures in logic programs: derivation by means of abstract interpretation, volume 675 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-56694-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 220 pp. LCCN QA76.9.D35 M84 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0675.
htm;   http://www.springerlink.
com/content/978-0-387-
56694-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=675.

Muller:1996:DIA

[Mul96] J. P. (Jorg P.) Muller.

The design of intelligent agents: a layered approach, volume 1177 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-62003-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 227 pp. LCCN TJ217.5.M84 1996. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1177.
htm;   http://www.springerlink.
com/content/978-3-540-
62003-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1177.

Muller:2002:MSV

[Mül02] Peter Müller. Modular Specification and Verification of Object-Oriented Programs, volume 2262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43167-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 292 pp. LCCN QA76.64 .M85 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2262.htm;
 http://www.springerlink.

com/content/978-3-540-
43167-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2262.

Based on the author’s thesis (doctoral)–Fernuniversität Hagen, 2001.

Muller:2007:SCF

[Mül07a] Christian Müller, editor. Speaker Classification I: Fundamentals, Features, and Methods, volume 4343 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74186-0 (print), 3-540-74200-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74200-5.

Muller:2007:SCI

[Mül07b] Christian Müller, editor. Speaker Classification II: Selected Projects, volume 4441 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74121-6 (print), 3-540-74122-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74122-0.

Muller:2010:ALS

[Mül10] Peter Müller, editor. Advanced Lectures on Software Engineering: LASER Summer School 2007/2008, volume 6029 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13009-7 (print), 3-642-13010-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13010-6.

Murray:1999:ARA

[Mur99] Niel V. Murray, editor. Automated reasoning with analytic tableaux and related methods: International Conference, TABLEAUX ’99, Saratoga Springs, NY, USA, June 7–11, 1999: Proceedings, volume 1617 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66086-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.9.A96 T33 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1617.htm;
http://www.springerlink.
com/content/978-3-540-
66086-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1617.

Menezes:1991:ACC

[MV91] Alfred J. Menezes and Scott A. Vanstone, editors. Advances in cryptology — CRYPTO ’90: proceedings, volume 537 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54508-5 (New York), 3-540-54508-5 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0537.htm;
http://www.springerlink.
com/content/978-0-387-
54508-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=537.

Conference held Aug. 11– 15, 1990, at the University of California, Santa Barbara.

Munch:2006:PFS

[MV06] Jürgen Münch and Matias Vierimaa, editors. Product-Focused Software Process Improvement: 7th International Conference, PROFES 2006, Amsterdam, The Netherlands, June 12– 14, 2006. Proceedings, volume 4034 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34682-1 (print), 3-540-34683-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34683-8.

Murphy:2007:CML

[MV07] Amy L. Murphy and Jan Vitek, editors. Coordination Models and Languages: 9th International Conference, COORDI-NATION 2007, Paphos, Cyprus, June 6–8, 2007. Proceedings, volume 4467 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72793-0 (print), 3-540-72794-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72794-1.

Miri:2012:SAC

[MV12a] Ali Miri and Serge Vaudenay, editors. Selected Areas in Cryptography: 18th International Workshop, SAC 2011, Toronto, ON, Canada, August 11–12, Revised Selected Papers, volume 7118 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28495-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28495-3.

Mitrokotsa:2012:PCA

[MV12b] Aikaterini Mitrokotsa and Serge Vaudenay, editors. Progress in Cryptology — AFRICACRYPT 2012: 5th International Conference on Cryptology in Africa, Ifrance, Morocco, July 10–12, 2012. Proceedings, volume 7374 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31409-0 (print), 3-642-31410-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31410-0.

Monostori:2001:EIS

[MVA01] László Monostori, József Váncza, and Moonis Ali, editors. Engineering of intelligent systems: 14th International Conference on Industrial and Engineering Applications of Artificial Intelligence and Expert Systems, IEA/AIE 2001, Budapest, Hungary, June 4–7, 2001: Proceedings, volume 2070 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42219-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I544 2001; QA267.A1 L43 no.2070. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2070.htm;
 http://www.springerlink.
com/content/978-3-540-
42219-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2070.

Marik:2007:HMA

[MVC07] Vladimír Mařík, Valeriy Vyatkin, and Armando W. Colombo, editors. Holonic and Multi-Agent Systems for Manufacturing: Third International Conference on Industrial Applications of Holonic and Multi-Agent Systems, Holo-MAS 2007, Regensburg, Germany, September 3– 5, 2007, volume 4659 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74478-9 (print), 3-540-74481-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74481-8.

Marron:2010:RWW

[MVCM10] Pedro J. Marron, Thiemo Voigt, Peter Corke, and Luca Mottola, editors. Real-World Wireless Sensor Networks: 4th International Workshop, REALWSN 2010, Colombo, Sri Lanka, December 16–17, 2010. Proceedings, volume 6511 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17519-8 (print), 3-642-17520-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17520-6.

Monteiro:2010:ICS

[MvL10] José Monteiro and René van Leuken, editors. Integrated Circuit and System Design. Power and Timing Modeling, Optimization and Simulation: 19th International Workshop, PATMOS 2009, Delft, The Netherlands, September 9–11, 2009, Revised Selected Papers, volume 5953 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11801-1 (print), 3-642-11802-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11802-9.

Marik:2011:HMA

[MVL11] Vladimír Mařík, Pavel Vrba, and Paulo Leitão, editors. Holonic and Multi-Agent Systems for Manufacturing: 5th International Conference on Industrial Applications of Holonic and Multi-Agent Systems, HoloMAS 2011, Toulouse, France, August 29–31, 2011. Proceedings, volume 6867 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23180-2 (print), 3-642-23181-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23181-0.

Martan-Vide:2004:MCI

[MVMP^{+}04] Carlos Martàn-Vide, Giancarlo Mauri, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa, editors. Membrane Computing: International Workshop, WMC 2003: Tarragona, Spain, July 17–22, 2003: Revised Papers, volume 2933 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20895-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.887 .W48 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2933.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20895-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2933;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95207.

Monien:1986:SAS

[MVN86] B. Monien and G. Vidal-Naquet, editors. STACS 86: 3rd Annual Symposium on Theoretical Aspects of Computer Science, Orsay, France, January 16–18, 1986, volume 210 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CODEN LNCSD9. ISBN 0-387-16078-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.210. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0210.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        16078-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=210.

Organized jointly by the special interest group for theoretical computer science of the Gesellschaft für Informatik (G.I.) and the special interest group for applied mathematic[s] of the Association française des sciences et techniques de l’information, de l’organisation et des systèmes (AFCET)”.

Martin-Vide:2008:LAT

[MVOF08] Carlos Martín-Vide, Friedrich Otto, and Henning Fernau, editors. Language and Automata Theory and Applications: Second International Conference, LATA 2008, Tarragona, Spain, March 13–19, 2008. Revised Papers, volume 5196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88281-2 (print), 3-540-88282-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88282-4.

Middeldorp:2005:PTC

[MvOvRdV05] Aart Middeldorp, Vincent van Oostrom, Femke van Raamsdonk, and Roel de Vrijer, editors. Processes, terms and cycles: steps on the road to infinity: essays dedicated to Jan Willem Klop on the occasion of his 60th birthday, volume 3838 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30911-X. ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 638 pp. LCCN QA267 .P76 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30911-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3838.

Margenov:2009:NAA

[MVW09] Svetozar Margenov, Lubin G. Vulkov, and Jerzy Waśniewski, editors. Numerical Analysis and Its Applications: 4th International Conference, NAA 2008, Lozenetz, Bulgaria, June 16–20, 2008. Revised Selected Papers, volume 5434 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00463-6 (print), 3-642-00464-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00464-3.

Maluszynski:1991:PLI

[MW91] J. Małuszynski and M. Wirsing, editors. Programming language implementation and logic programming: 3rd International Symposium, PLILP ’91, Passau, Germany, August 26– 28, 1991: proceedings, volume 528 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54444-5 (Berlin), 0-387-54444-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .I5773 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0528.htm;
http://www.springerlink.
com/content/978-3-540-
54444-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=528.

Mazurkiewicz:1997:CCT

[MW97] Antoni Mazurkiewicz and Józef Winkowski, editors. CONCUR ’97: concurrency theory: 8th International Conference, Warsaw, Poland, July 1–4, 1997: proceedings, volume 1243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63141-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1243.htm;
 http://www.springerlink.
com/content/978-3-540-
63141-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1243.

Maurer:2003:EPA

[MW03] Frank Maurer and Don Wells, editors. Extreme Programming and Agile Methods—XP/Agile Universe 2003: third XP Universe and second Agile Universe Conference, New Orleans, LA, USA, August 10–13, 2003: Proceedings, volume 2753 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40662-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 X63 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2753.
htm;   http://www.springerlink.
com/content/978-3-540-
40662-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2753;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11928.

Meyer:2008:VST

[MW08] Bertrand Meyer and Jim Woodcock, editors. Verified Software: Theories, Tools, Experiments: First IFIP TC 2/WG 2.3 Conference, VSTTE 2005, Zurich, Switzerland, October 10–13, 2005, Revised Selected Papers and Discussions, volume 4171 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69147-2 (print), 3-540-69149-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69149-5.

McGinty:2009:CBR

[MW09] Lorraine McGinty and David C. Wilson, editors. Case-Based Reasoning Research and Development: 8th International Conference on Case-Based Reasoning, IC-CBR 2009 Seattle, WA, USA, July 20–23, 2009 Proceedings, volume 5650 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02997-3 (print), 3-642-02998-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02998-1.

Marron:2011:WSN

[MW11] Pedro José Marrón and Kamin Whitehouse, editors. Wireless Sensor Networks: 8th European Conference, EWSN 2011, Bonn, Germany, February 23–25, 2011. Proceedings, volume 6567 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19185-1 (print), 3-642-19186-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19186-2.

MacCaull:2006:RMC

[MWD06] Wendy MacCaull, Michael Winter, and Ivo Düntsch, editors. Relational Methods in Computer Science: 8th International Seminar on Relational Methods in Computer Science, 3rd International Workshop on Applications of Kleene Algebra, and Workshop of Cost Action 274: Tarski, St. Catharines, ON, Canada, February 22–26, 2005, Selected Revised Papers, volume 3929 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33339-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33339-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3929.

Muller:1997:IAI

[MWJ97] J. P. Muller, Michael J.

Wooldridge, and Nick Jennings, editors. Intelligent agents III: agent theories, architectures, and languages: ICAI’96 Workshop (ATAL), Budapest, Hungary, August 12–13, 1996: proceedings, volume 1193 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62507-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1193. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1193.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62507-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1193.

Muneesawang:2009:AMI

[MWK^{+}09] Paisarn Muneesawang, Feng Wu, Itsuo Kumazawa, Athikom Roeksabutr, Mark Liao, and Xiaoou Tang, editors. Advances in Multimedia Information Processing — PCM 2009: 10th Pacific Rim Conference on Multimedia, Bangkok, Thailand, December 15–18, 2009 Proceedings, volume 5879 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10466-5 (print), 3-642-10467-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10467-1.

Margenov:2001:LSS

[MWY01] Svetozar Margenov, Jerzy Waśniewski, and Plamen Yalamov, editors. Largescale scientific computing: Third International Conference, LSSC 2001, Sozopol, Bulgaria, June 6–10, 2001: Revised Papers, volume 2179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43043-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .L77 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2179.htm;
 http://www.springerlink.
com/content/978-3-540-
43043-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2179.

Maglio:2010:SOC

[MWYF10] Paul P. Maglio, Mathias Weske, Jian Yang, and Marcelo Fantinato, editors. Service-Oriented Computing: 8th International Conference, IC-SOC 2010, San Francisco, CA, USA, December 7–10, 2010. Proceedings, volume 6470 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17357-8 (print), 3-642-17358-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17358-5.

Massacci:2010:ESS

[MWZ10] Fabio Massacci, Dan Wallach, and Nicola Zannone, editors. Engineering Secure Software and Systems: Second International Symposium, ES-SoS 2010, Pisa, Italy, February 3–4, 2010. Proceedings, volume 5965 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11746-5 (print), 3-642-11747-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11747-3.

Megiddo:2005:AAM

[MXZ05] Nimrod Megiddo, Yinfeng Xu, and Binhai Zhu, editors. Algorithmic applications in management: first international conference, AAIM 2005, Xian, China, June 22–25, 2005, proceedings, volume 3521 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26224-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A25 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26224-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3521;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b137095.

Margaria:2001:TAC

[MY01] Tiziana Margaria and Wang Yi, editors. Tools and algorithms for the construction and analysis of systems: 7th International Conference, TACAS 2001, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2001, Genova, Italy, April 2– 6, 2001: Proceedings, volume 2031 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41865-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 T33 2001; QA267.A1 L43 no.2031. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2031.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41865-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2031.

Mao:2006:DCI

[MY06] Chengde Mao and Takashi Yokomori, editors. DNA Computing: 12th International Meeting on DNA Computing, DNA12, Seoul, Korea, June 5– 9, 2006, Revised Selected Papers, volume 4287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49024-8 (print), 3-540-68423-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68423-7.

Mycroft:1995:SAS

[Myc95] Alan Mycroft, editor. Static analysis: Second International Symposium, SAS ’95, Glasgow, UK, September 25–27, 1995: proceedings, volume 983 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60360-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0983.htm;
 http://www.springerlink.
com/content/978-3-540-
60360-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=983.

Munch:2010:NMC

[MYS10] Jürgen Münch, Ye Yang, and Wilhelm Schäfer, editors. New Modeling Concepts for Today’s Software Processes: International Conference on Software Process, ICSP 2010, Paderborn, Germany, July 8–9, 2010. Proceedings, volume 6195 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14346-6 (print), 3-642-14347-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14347-2.

McCurley:1998:ACE

[MZ98] Kevin S. McCurley and Claus Dieter Ziegler, editors. Advances in Cryptology, 1981–1997: Electronic Proceedings and Index of the CRYPTO and EUROCRYPT Conferences, 1981–1997, volume 1440 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65069-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 A373 1998.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1440.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65069-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1440.

CD-ROM contains 14692 pages of information from 32 volumes of conference proceedings of CRYPTO and EURO-CRYPT.

Mambo:1999:ISS

[MZ99] Masahiro Mambo and Yuliang Zheng, editors. Information security: Second International Workshop, ISW’99, Kuala Lumpur, Malaysia, November 6–7, 1999:

Proceedings, volume 1729 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66695-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1729. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1729.htm;
 http://www.springerlink.
com/content/978-3-540-
66695-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1729.

Matsui:2004:SAC

[MZ04] Mitsuru Matsui and Robert Zuccherato, editors. Selected Areas in Cryptography: 10th Annual International Workshop, SAC 2003, Ottawa, Canada, August 14–15, 2003, Revised Papers, volume 3006 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21370-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3006.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        21370-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3006;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96837.

Muller:2006:AOS

[MZ06a] Jörg P. Müller and Franco Zambonelli, editors. Agent-Oriented Software Engineering VI: 6th International Workshop, AOSE 2005, Utrecht, The Netherlands, July 25, 2005. Revised and Invited Papers, volume 3950 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34097-1 (print), 3-540-34099-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34099-7.

Mycroft:2006:CCI

[MZ06b] Alan Mycroft and Andreas Zeller, editors. Compiler Construction: 15th International Conference, CC 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 30–31, 2006. Proceedings, volume 3923 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33050-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33050-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3923.

Ma:2007:CPM

[MZ07a] Bin Ma and Kaizhong Zhang, editors. Combinatorial Pattern Matching: 18th Annual Symposium, CPM 2007, London, Canada, July 9– 11, 2007. Proceedings, volume 4580 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73436-8 (print), 3-540-73437-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73437-6.

Mandoiu:2007:BRA

[MZ07b] Ion Măndoiu and Alexander Zelikovsky, editors. Bioinformatics Research and Applications: Third International Symposium, ISBRA 2007, Atlanta, GA, USA, May 7–10, 2007. Proceedings, volume 4463 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72030-8 (print), 3-540-72031-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72031-7.

Mehrotra:2006:ISI

[MZC06] Sharad Mehrotra, Daniel D. Zeng, and Hsinchun Chen, editors. Intelligence and Security Informatics: IEEE International Conference on Intelligence and Security Informatics, ISI 2006, San Diego, CA, USA, May 23–24, 2006. Proceedings, volume 3975 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34478-0 (softcover). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
34478-0.

Mauthe:2009:FMN

[MZCC09] Andreas Mauthe, Sherali Zeadally, Eduardo Cerqueira, and Marília Curado, editors. Future Multimedia Networking: Second International Workshop, FMN 2009, Coimbra, Portugal, June 22–23, 2009. Proceedings, volume 5630 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02471-8 (print), 3-642-02472-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02472-6.

Mundy:1994:AIC

[MZF94] Joseph L. Mundy, Andrew Zisserman, and David Forsyth, editors. Applications of invariance in computer vision: second joint European-US workshop, Ponta Delgada, Azores, Portugal, October 9–14, 1993: proceedings, volume 825 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58240-1 (Berlin), 0-387-58240-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .A65 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0825.htm;
 http://www.springerlink.
com/content/978-3-540-
58240-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=825.

“Proceedings of the Second Joint European-US Workshop on ‘Applications of Invariance in Computer Vision”’–Pref.

Myaeng:2005:IRT

[MZWZ05] Sung Hyon Myaeng, Ming Zhou, Kam-Fai Wong, and Hong-Jiang Zhang, editors. Information retrieval technology: Asia Information Retrieval Symposium, AIRS 2004, Beijing, China, October 18–20, 2004: Revised selected papers, volume 3411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25065-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z699.A1 A75 2004. URL

http://springerlink.

metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3411;
http://www.springerlink.
com/content/978-3-540-
25065-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3411;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106653.

Nwana:1997:SAS

[NA97] Hyacinth Sama Nwana and Nader Azarmi. Software agents and soft computing: towards enhancing machine intelligence: concepts and applications, volume 1198 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62560-7. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 298 pp. LCCN QA76.76.I58 S642 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1198.htm;
http://www.springerlink.
com/content/978-3-540-
62560-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1198.

Noda:2012:SMP

[NABK12] Itsuki Noda, Noriaki Ando, Davide Brugali, and James J. Kuffner, editors. Simulation, Modeling, and Programming for Autonomous Robots: Third International Conference, SIMPAR 2012, Tsukuba, Japan, November 5–8, 2012. Proceedings, volume 7628 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34326-0 (print), 3-642-34327-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34327-8.

Naccache:2001:TCC

[Nac01] David Naccache, editor. Topics in cryptology, CT-RSA 2001: the Cryptographers’ Track at RSA Conference 2001, San Francisco, CA, USA, April 8–12, 2001: Proceedings, volume 2020 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41898-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA267.A1 L43 no.2020; QA76.9.A25 R753 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2020.htm;
http://www.springerlink.
com/content/978-3-540-
41898-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2020.

Naccache:2012:CST

[Nac12] David Naccache, editor.

Cryptography and Security: From Theory to Applications: Essays Dedicated to Jean-Jacques Quisquater on the Occasion of His 65th Birthday, volume 6805 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28367-5. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-28367-3.

Nadathur:1999:PPD

[Nad99] Gopalan Nadathur, editor. Principles and practice of declarative programming: International Conference PPDP’99, Paris, France, September 29–October 1, 1999: Proceedings, volume 1702 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66540-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.615 .P57 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1702.htm;
 http://www.springerlink.
com/content/978-3-540-
66540-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1702.

Nagl:1990:GTC

[Nag90] Manfred Nagl, editor. Graph-theoretic concepts in computer science: 15th International Workshop WG ’89, Castle Rolduc, The Netherlands, June 14–16, 1989: proceedings, volume 411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52292-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G7172 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0411.htm;
 http://www.springerlink.
com/content/978-0-387-
52292-0;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=411.

Revised versions of nearly all the papers presented at the 15th International Workshop on Graph-Theoretic Concepts in Computer Science (WG ’89) organized by Lehrstuhl für Informatic III of Aachen University of Technology.

Nagl:1995:GTC

[Nag95] Manfred Nagl, editor.

Graph-theoretic concepts in computer science: 21st International Workshop, WG ’95, Aachen, Germany, June 20–22, 1995: proceedings, volume 1017 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60618-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .I572 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1017.htm;
http://www.springerlink.
com/content/978-3-540-
60618-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1017.

Nagl:1996:BTI

[Nag96] Manfred Nagl, editor.

Building tightly integrated software development environments: the IPSEN approach, volume 1170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61985-2. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 709 pp. LCCN QA76.76.O63B838 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1170.
htm;   http://www.springerlink.
com/content/978-3-540-
61985-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1170.

Nagi:2001:TAT

[Nag01] Khaled Nagi. Transactional Agents: Towards a Robust Multi-Agent System, volume 2249 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43046-6. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 205 pp. LCCN QA76.76.I58 N34 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2249.htm;
 http://www.springerlink.
com/content/978-3-540-

43046-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2249.

Naish:1986:NCP

[Nai86] Lee Naish. Negation and control in PRO-LOG, volume 238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16815-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 119 pp. LCCN QA76.73.P76 N351 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0238.htm;
http://www.springerlink.
com/content/978-0-387-
16815-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=238.

Nakamura:1992:PIA

[Nak92] A. Nakamura, editor.

Parallel image analysis: second international conference, ICPIA ’92, Ube, Japan, December 21–23, 1992: proceedings, volume 654 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-56346-6 (Berlin), 0-387-56346-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .P375 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0654.htm;
 http://www.springerlink.
com/content/978-3-540-
56346-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=654.

Naor:2004:TCF

[Nao04] Moni Naor, editor. Theory of Cryptography: First Theory of Cryptography Conference, TCC 2004, Cambridge, MA, USA, February 19–21, 2004: Proceedings, volume 2951 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21000-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C6676 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2951.
htm;   http://www.springerlink.
com/content/978-3-540-
21000-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2951;
 http://www.springerlink.
com/openurl.asp?genre=

volume&id=doi:10.1007/
b95566.

Naor:2007:ACE

[Nao07] Moni Naor, editor. Advances in Cryptology — EUROCRYPT 2007: 26th Annual International Conference on the Theory and Applications of Cryptographic Techniques, Barcelona, Spain, May 20–24, 2007. Proceedings, volume 4515 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-72540-7. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72540-4.

Nareyek:2001:CBA

[Nar01a] Alexander Nareyek. Constraint-Based Agents: An Architecture for Constraint-Based Modeling and Local-Search-Based Reasoning for Planning and Scheduling in Open and Dynamic Worlds, volume 2062 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42258-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 178 pp. LCCN QA76.73.I58 N37 2001; QA267.A1 L43 no.2062. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2062.htm;
 http://www.springerlink.
com/content/978-3-540-
42258-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2062.

Nareyek:2001:LSP

[Nar01b] Alexander Nareyek, editor. Local search for planning and scheduling: ECAI 2000 Workshop, Berlin, Germany, August 21, 2000: Revised Papers, volume 2148 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42898-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.I58 W67 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2148.htm;
 http://www.springerlink.
com/content/978-3-540-
42898-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2148.

Based on contributions to the Workshop on Local Search for Planning and Scheduling, held during ECAI 2000.

Naumann:2002:QDQ

[Nau02] Felix Naumann. Quality-Driven Query Answering for Integrated Information Systems, volume 2261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43349-X. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 166 pp. LCCN QA76.9.W43 N38 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2261.htm;
 http://www.springerlink.
com/content/978-3-540-
43349-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2261.

Nayak:1995:AMP

[Nay95] P. Pandurang Nayak.

Automated modeling of physical systems, volume 1003 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60641-6 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). xx + 232 pp. LCCN QA76.9.C65N39 1995.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1003.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60641-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1003.

Nazareth:1994:NCF

[Naz94] J. L. (John L.) Nazareth. The Newton–Cauchy framework: a unified approach to unconstrained nonlinear minimization, volume 769 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-57671-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 101 pp. LCCN T57.8 .N39 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0769.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57671-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=769.

Nolfi:2006:AAI

[NBC^{+}06] Stefano Nolfi, Gianluca Baldassarre, Raffaele Calabretta, John C. T. Hallam, Davide Marocco, Jean-Arcady Meyer, Orazio Miglino, and Domenico Parisi, editors. From Animals to Animats 9: 9th International Conference on Simulation of Adaptive Behavior, SAB 2006, Rome, Italy, September 25–29, 2006. Proceedings, volume 4095 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38608-4 (print), 3-540-38615-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38615-5.

Nguyen:2008:NFA

[NBGA08] Ngoc Thanh Nguyen, Leszek Borzemski, Adam Grzech, and Moonis Ali, editors. New Frontiers in Applied Artificial Intelligence: 21st International Conference on Industrial, Engineering and Other Applications of Applied Intelligent Systems, IEA/AIE 2008 Wrocław, Poland, June 18–20, 2008 Proceedings, volume 5027 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69045-X (print), 3-540-69052-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69052-8.

Nikolaidis:2004:AHM

[NBK04] Ioanis Nikolaidis, Michel Barbeau, and Evangelos Kranakis, editors. Ad-Hoc, Mobile, and Wireless Networks: Third International Conference, ADHOC-NOW 2004, Vancouver, Canada, July 22–24, 2004: Proceedings, volume 3158 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22543-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK6570.M6 A34 2004.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3158.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        22543-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3158;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=

volume&id=doi:10.1007/
b99253.

Nunez:2009:TSC

[NBM09] Manuel Núñez, Paul Baker, and Mercedes G. Merayo, editors. Testing of Software and Communication Systems: 21st IFIP WG 6.1 International Conference, TEST-COM 2009 and 9th International Workshop, FATES 2009, Eindhoven, The Netherlands, November 2–4, 2009. Proceedings, volume 5826 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05030-1 (print), 3-642-05031-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
05031-2.

Nicosia:2004:AIS

[NCBT04] Giuseppe Nicosia, Vincenzo Cutello, Peter J. Bentley, and Jon Timmis, editors. Artificial Immune Systems: Third International Conference, ICARIS 2004, Catania, Sicily, Italy, September 13–16, 2004, Proceedings, volume 3239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23097-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3239.htm;
 http://www.springerlink.
com/content/978-3-540-
23097-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3239;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100486.

Nakashima:1999:AIA

[NcC99] Hideyuki Nakashima and Ch’eng chi Chang, editors. Approaches to intelligent agents: second Pacific Rim International Workshop on Multi-Agents, PRIMA’99, Kyoto, Japan, December 2–3, 1999: Proceedings, volume 1733 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66823-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1733.

URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t1733.
htm;   http://www.springerlink.
com/content/978-3-540-
66823-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1733.

Nikoletseas:2008:DCS

[NCJK08] Sotiris E. Nikoletseas, Bogdan S. Chlebus, David B. Johnson, and Bhaskar Krishnamachari, editors. Distributed Computing in Sensor Systems: 4th IEEE International Conference, DCOSS 2008 Santorini Island, Greece, June 11– 14, 2008 Proceedings, volume 5067 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69169-3 (print), 3-540-69170-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69170-9.

Nienhuys-Cheng:1997:FIL

[NCW97] S.-H. (Shan-Hwei) Nienhuys-Cheng and R. Wolf. Foundations of inductive logic programming, volume 1228 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62927-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xvii + 404 pp. LCCN QA76.63 .N54 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1228.
htm;   http://www.springerlink.
com/content/978-3-540-
62927-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1228.

Nurnberger:2004:AMR

[ND04] Andreas Nürnberger and Marcin Detyniecki, editors. Adaptive Multimedia Retrieval: First International Workshop, AMR 2003, Hamburg, Germany, September 15– 16, 2003, Revised Selected and Invited Papers, volume 3094 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22163-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3094.htm;
 http://www.springerlink.

com/content/978-3-540-
22163-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3094;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98244.

Natkin:2009:ECI

[ND09] Stéphane Natkin and Jérôme Dupire, editors. Entertainment Computing — ICEC 2009: 8th International Conference, Paris, France, September 3–5, 2009. Proceedings, volume 5709 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04051-9 (print), 3-642-04052-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04052-8.

Naftalin:1994:FIB

[NDB94] Maurice Naftalin, Tim Denvir, and Miquel Bertran, editors. FME ’94: industrial benefit of formal methods: Second International Symposium of Formal Methods Europe, Barcelona, Spain, October 24–28, 1994: proceedings, volume 873 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58555-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I593 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0873.
htm;   http://www.springerlink.
com/content/978-0-387-
58555-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=873.

Nebel:1994:KAA

[NDF94] Bernhard Nebel and Leonie Dreschler-Fischer, editors. KI-94: advances in artificial intelligence: 18th German Annual Conference on Artificial Intelligence, Saarbrucken, Germany, September 18–23, 1994: proceedings, volume 861 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58467-6 (Berlin), 0-387-58467-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G466 1994.

URL http://link.springer-
ny.com/link/service/

series/0558/tocs/t0861.
htm;   http://www.springerlink.
com/content/978-3-540-
58467-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=861.

Nagl:1997:GTC

[NDFMS97] Manfred Nagl, Fabrizio D’Amore, Paolo G. Franciosa, and Alberto Marchetti-Spaccamela, editors. Graph-theoretic concepts in computer science: 22nd International Workshop, WG ’96, Cadenabbia, Italy, June 12–14, 1996: proceedings, volume 1197 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62559-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1197.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1197.
htm;   http://www.springerlink.
com/content/978-3-540-
62559-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1197.

Niemegeers:2004:PWC

[NdG04] Ignas Niemegeers and Sonia Heemstra de Groot, editors. Personal Wireless Communications:

IFIP TC6 9th International Conference, PWC 2004, Delft, The Netherlands, September 21– 23, 2004, Proceedings, volume 3260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23162-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3260.htm;
 http://www.springerlink.
com/content/978-3-540-
23162-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3260;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100819.

Nascimento:2003:SPI

[NdMO03] Mario A. Nascimento, Edleno S. de Moura, and Arlindo L. Oliveira, editors. String Processing and Information Retrieval: 10th International Symposium, SPIRE 2003, Manaus, Brazil, October 8–10, 2003: Proceedings, volume 2857 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20177-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.T48 S68 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2857.
htm;   http://www.springerlink.
com/content/978-3-540-
20177-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2857;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b14038.

Nielsen:2002:FSS

[NE02] M. Nielsen and Uffe Engberg, editors. Foundations of software science and computation structures: 5th International Conference, FOS-SACS 2002, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2002, Grenoble, France, April 8–12, 2002: Proceedings, volume 2303 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43366-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F67 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2303.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43366-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2303.

Nebel:1990:RRH

[Neb90] Bernhard Nebel. Reasoning and revision in hybrid representation systems, volume 422 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52443-6 (Berlin), 0-387-52443-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 270 pp. LCCN Q335 .N43 1990. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0422.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        52443-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=422.

Revision of the author’s thesis (Saarland, 1989).

Nehmer:1988:EDS

[Neh88] J. Nehmer, editor. Experiences with distributed systems: international workshop, Kaiserslautern, FRG, September 28– 30, 1987: proceedings, volume 309 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19333-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 E871 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0309.htm;
 http://www.springerlink.
com/content/978-0-387-
19333-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=309.

Sponsored and organized by the Sonderforschungsbereich 124 “VLSI-Entwurfsmethoden und Parallelität” and the ENC of the IBM Corporation, Heidelberg.

Nehaniv:1999:CMA

[Neh99] Chrystopher L. Nehaniv, editor. Computation for Metaphors, Analogy, and Agents, volume 1562 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65959-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 388 pp. LCCN Q335 .C555 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1562.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65959-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1562.

Nesi:1995:OSQ

[Nes95] Paolo Nesi, editor. Objective software quality: objective quality: Second Symposium on Software Quality Techniques and Acquisition Criteria, Florence, Italy, May 28–30, 1995: proceedings, volume 926 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59449-3, 0-387-59449-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.Q35 S96 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0926.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        59449-9;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=926.

Nesetril:1999:AEA

[Nes99] Jaroslav Nesetril, editor. Algorithms — ESA ’99: 7th annual European symposium, Prague, Czech Republic, July 16–18, 1999: Proceedings, volume 1643 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66251-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.93 E83 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1643.htm;
 http://www.springerlink.
com/content/978-3-540-
66251-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1643.

Neuenschwander:2004:PSM

[Neu04] Daniel Neuenschwander.

Probabilistic and Statistical Methods in Cryptology: An Introduction by Selected Topics, volume 3028 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22001-1. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 158 pp. LCCN QA76 A1 L43 3028. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3028.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22001-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3028;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97045.

Nielson:2007:SAI

[NF07] Hanne Riis Nielson and Gilberto Filé, editors. Static Analysis: 14th International Symposium, SAS 2007, Kongens Lyngby, Denmark, August 22–24, 2007. Proceedings, volume 4634 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74060-0 (print), 3-540-74061-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74061-2.

Ng:1979:SAC

[Ng79] Edward W. Ng, editor. Symbolic and algebraic computation: EU-ROSAM ’79, an International Symposium on Symbolic and Algebraic Manipulation, Marseille, France, June 1979, volume 72 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CODEN LNCSD9. ISBN 0-387-09519-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E88 1979.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0072.
htm;   http://www.springerlink.
com/content/978-0-387-
09519-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=72.

Ng:1993:ACI

[Ng93] K. W. Ng, editor. Algorithms and computation: 4th International Symposium, ISAAC ’93, Hong Kong, December 1993: proceedings, volume 762 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57568-5 (Berlin), 0-387-57568-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I73 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0762.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        57568-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=762.

Norrie:2010:ODS

[NG10] Moira C. Norrie and Michael Grossniklaus, editors. Object Databases: Second International Conference, ICOODB 2009, Zurich, Switzerland, July 1–3, 2009. Revised Papers, volume 5936 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14680-5 (print), 3-642-14681-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14681-7.

Nguyen:2007:AMA

[NGHJ07] Ngoc Thanh Nguyen, Adam Grzech, Robert J. Howlett, and Lakhmi C. Jain, editors. Agent and Multi-Agent Systems: Technologies and Applications: First KES International Symposium, KES-AMSTA 2007, Wroclaw, Poland, May 31– June 1, 2007. Proceedings, volume 4496 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72829-5 (print), 3-540-72830-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72830-6.

Ngo:2009:CCA

[Ngo09] Hung Q. Ngo, editor.

Computing and Combinatorics: 15th Annual International Conference, COCOON 2009 Niagara Falls, NY, USA, July 13–15, 2009 Proceedings, volume 5609 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02881-0 (print), 3-642-02882-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02882-3.

Nguyen:2006:PCV

[Ngu06] Phong Q. Nguyen, editor. Progress in Cryptology — VIETCRYPT 2006: First International Conference on Cryptology in Vietnam, Hanoi, Vietnam, September 25– 28, 2006. Revised Selected Papers, volume 4341 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68799-8 (print), 3-540-68800-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68800-6.

Nguyen:2011:TCCa

[Ngu11a] Ngoc Thanh Nguyen, editor. Transactions on Computational Collective Intelligence III, volume 6560 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19967-4 (print), 3-642-19968-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19968-4.

Nguyen:2011:TCCb

[Ngu11b] Ngoc Thanh Nguyen, editor. Transactions on Computational Collective Intelligence IV, volume 6660 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21883-0 (print), 3-642-21884-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21884-2.

Nguyen:2011:TCCc

[Ngu11c] Ngoc Thanh Nguyen, editor. Transactions on Computational Collective Intelligence V, volume 6910 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24015-1 (print), 3-642-24016-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24016-4.

Nguyen:2012:TCCa

[Ngu12a] Ngoc Thanh Nguyen, editor. Transactions on Computational Collective Intelligence VI, volume 7190 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29355-7 (print), 3-642-29356-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29356-6.

Nguyen:2012:TCCb

[Ngu12b] Ngoc Thanh Nguyen, editor. Transactions on Computational Collective Intelligence VII, volume 7270 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32065-1 (print), 3-642-32066-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32066-8.

Nguyen:2012:TCCc

[Ngu12c] Ngoc-Thanh Nguyen, editor. Transactions on Computational Collective Intelligence VIII, volume 7430 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34644-8 (print), 3-642-34645-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34645-3.

Nyberg:2003:SAC

[NH03] Kaisa Nyberg and Howard Heys, editors. Selected areas in cryptography: 9th annual International Workshop, SAC 2002, St. John’s, Newfoundland, Canada, August 15–16, 2002: Revised Papers, volume 2595 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00622-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 S22 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2595.
htm;   http://www.springerlink.
com/content/978-3-540-
00622-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2595.

Also available via the World Wide Web.

Negoita:2004:KBIa

[NHJ04a] Mircea Gh. Negoita, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 8th International Conference, KES 2004, Wellington, New Zealand, September 20– 25, 2004, Proceedings, Part I, volume 3213 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23318-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3213.htm;
 http://www.springerlink.
com/content/978-3-540-
23318-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3213;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100909.

Negoita:2004:KBIb

[NHJ04b] Mircea Gh. Negoita, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 8th International Conference, KES 2004, Wellington, New Zealand, September 20– 25, 2004, Proceedings, Part II, volume 3214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23206-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3214.htm;
 http://www.springerlink.
com/content/978-3-540-
23206-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3214;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100910.

Negoita:2004:KBIc

[NHJ04c] Mircea Gh. Negoita, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 8th International Conference, KES 2004, Wellington, New Zealand, September 20– 25, 2004, Proceedings, Part III, volume 3215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23205-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3215.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23205-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3215;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100916.

Nguyen:2012:CCIa

[NHJ12a] Ngoc-Thanh Nguyen, Kiem Hoang, and Piotr Jȩdrzejowicz, editors. Computational Collective Intelligence. Technologies and Applications: 4th International Conference, ICCCI 2012, Ho Chi Minh City, Vietnam, November 28–30, 2012, Proceedings, Part I, volume 7653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34629-4 (print), 3-642-34630-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34630-9.

Nguyen:2012:CCIb

[NHJ12b] Ngoc-Thanh Nguyen, Kiem Hoang, and Piotr Jȩdrzejowicz, editors. Computational Collective Intelligence. Technologies and Applications: 4th International Conference, ICCCI 2012, Ho Chi Minh City, Vietnam, November 28–30, 2012, Proceedings, Part II, volume 7654 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34706-1 (print), 3-642-34707-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34707-8.

Nickel:1975:IMP

[Nic75] Karl Nickel, editor.

Interval mathematics: proceedings of the international symposium, Karlsruhe, West Germany, May 20–24, 1975, volume 29 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07170-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA297 .I541. DM30.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0029.

                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        07170-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=29.

German or English.

Nickel:1986:IMP

[Nic86] Karl Nickel, editor. Interval mathematics 1985: proceedings of the International Symposium, Freiburg i. Br., Federal Republic of Germany, September 23– 26, 1985, volume 212 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16437-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA297.75 .I581 1986. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0212.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        16437-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=212.

Nierstrasz:1993:EOO

[Nie93] Oscar Marius Nierstrasz, editor. ECOOP ’93, object-oriented programming, 7th European conference, Kaiserslautern, Germany, July 26–30, 1993: proceedings, volume 707 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57120-5 (Berlin), 0-387-57120-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0707.htm;
 http://www.springerlink.
com/content/978-3-540-
57120-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=707.

Nievergelt:1994:IGI

[Nie94] Jurg Nievergelt, editor.

IGIS ’94 — geographic information systems: International Workshop on Advanced Research in Geographic Information Systems, Monte Verita, Ascona, Switzerland, February/March, 1994: proceedings, volume 884 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-58795-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.2 .I62 1994. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0884.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        58795-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=884.

Nielson:1996:PLS

[Nie96] Hanne Riis Nielson, editor. Programming languages and systems, ESOP ’96: 6th European Symposium on Programming, Linköping, Sweden, April 22–24, 1996: proceedings, volume 1058 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61055-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1058.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61055-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1058.

Nielsen:1999:SST

[Nie99] Mads Nielsen, editor. Scale-space theories in computer vision: Second International Conference, Scale-Space’99, Corfu, Greece, September 26–27, 1999: Proceedings, volume 1682 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66498-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .S375 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1682.htm;
 http://www.springerlink.
com/content/978-3-540-
66498-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1682.

Nieuwenhuis:2003:RTA

[Nie03] Robert Nieuwenhuis, editor. Rewriting techniques and applications: 14th International Conference, RTA 2003, Valencia, Spain, June 2003: Proceedings, volume 2706 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40254-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I555 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2706.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        40254-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2706.

Nieuwenhuis:2005:ADC

[Nie05] Robert Nieuwenhuis, editor. Automated deduction, CADE-20: 20th International Conference on Automated Deduction, Tallinn, Estonia, July 22–27, 2005: proceedings, volume 3632 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28005-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I57 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28005-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3632.

Nielsen:2009:ETV

[Nie09] Frank Nielsen, editor. Emerging Trends in Visual Computing: LIX Fall Colloquium, ETVC 2008, Palaiseau, France, November 18–20, 2008. Revised Invited Papers, volume 5416 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00825-9 (print), 3-642-00826-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00826-9.

Nijholt:1980:CFG

[Nij80] Anton Nijholt. Contextfree grammars: covers, normal forms, and parsing, volume 93 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CODEN LNCSD9. ISBN 0-387-10245-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 253 pp. LCCN QA267.3 .N54. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0093.htm;
http://www.springerlink.
com/content/978-0-387-
10245-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=93.

Nikoletseas:2005:EEA

[Nik05] Sotiris E. Nikoletseas, editor. Experimental and efficient algorithms: 4th International Workshop, WEA 2005, Santorini Island, Greece, May 10–13, 2005, proceedings, volume 3503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25920-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25920-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3503;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136461.

Nipkow:1998:RTA

[Nip98] Tobias Nipkow, editor. Rewriting techniques and applications: 9th International Conference, RTA-98, Tsukuba, Japan, March 28–April 1, 1998: Proceedings, volume 1379 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64301-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.R48 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1379.

htm;   http://www.springerlink.
com/content/978-3-540-
64301-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1379.

Nivat:1998:FSS

[Niv98] Maurice Nivat, editor.

Foundations of software science and computation structures: First International Conference, FoSSaCS’98 held as part of the joint European Conferences on Theory and Practice of Software, ETAPS’98, Lisbon, Portugal, March/April 1998: Proceedings, volume 1378 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64300-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751.F67 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1378.htm;
http://www.springerlink.
com/content/978-3-540-
64300-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1378.

Noble:2009:TPL

[NJ09] James Noble and Ralph Johnson, editors. Transactions on Pattern Languages of Programming I, volume 5770 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10831-8 (print), 3-642-10832-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10832-7.

Navab:2010:IPC

[NJ10] Nassir Navab and Pierre Jannin, editors. Information Processing in Computer-Assisted Interventions: First International Conference, IPCAI 2010, Geneva, Switzerland, June 23, 2010. Proceedings, volume 6135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13710-5 (print), 3-642-13711-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13711-2.

Noble:2011:TPL

[NJA^{+}11] James Noble, Ralph Johnson, Paris Avgeriou, Neil B. Harrison, and Uwe Zdun, editors. Transactions on Pattern Languages of Programming II: Special Issue on Applying Patterns, volume 6510 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19431-1 (print), 3-642-19432-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19432-0.

Nguyen:2008:AMA

[NJHJ08] Ngoc Thanh Nguyen, Geun Sik Jo, Robert J. Howlett, and Lakhmi C. Jain, editors. Agent and Multi-Agent Systems: Technologies and Applications: Second KES International Symposium, KES-AMSTA 2008, Incheon, Korea, March 26–28, 2008. Proceedings, volume 4953 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78581-7 (print), 3-540-78582-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78582-8.

Nori:1988:FST

[NK88] Kesav V. Nori and S. Kumar, editors. Foundations of software technology and theoretical computer science: eighth conference, Pune, India, December 21–23, 1988: proceedings, volume 338 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-50517-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F681 1988. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0338.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        50517-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=338.

Nishio:1999:AMC

[NK99] Shojiro Nishio and Fumio Kishino, editors. Advanced multimedia content processing: First International Conference, AMCP ’98, Osaka, Japan, November 9–11, 1998: Proceedings, volume 1554 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65762-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .A52 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1554.htm;
http://www.springerlink.
com/content/978-3-540-
65762-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1554.

Nguyen:2010:TCCa

[NK10a] Ngoc Thanh Nguyen and Ryszard Kowalczyk, editors. Transactions on Computational Collective Intelligence I, volume 6220 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15033-0 (print), 3-642-15034-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15034-0.

Nguyen:2010:TCCb

[NK10b] Ngoc Thanh Nguyen and Ryszard Kowalczyk, editors. Transactions on Computational Collective Intelligence II, volume 6450 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17154-0 (print), 3-642-17155-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17155-0.

Nguyen:2009:CCI

[NKC09] Ngoc Thanh Nguyen, Ryszard Kowalczyk, and Shyi-Ming Chen, editors. Computational Collective Intelligence. Semantic Web, Social Networks and Multiagent Systems: First International Conference, ICCCI 2009, Wrocław, Poland, October 5–7, 2009. Proceedings, volume 5796 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04440-9 (print), 3-642-04441-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04441-0.

Nguyen:2011:IIDa

[NKJ11a] Ngoc Thanh Nguyen, Chong-Gun Kim, and Adam Janiak, editors. Intelligent Information and Database Systems: Third International Conference, ACIIDS 2011, Daegu, Korea, April 20– 22, 2011, Proceedings, Part I, volume 6591 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20038-9 (print), 3-642-20039-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20039-7.

Nguyen:2011:IIDb

[NKJ11b] Ngoc Thanh Nguyen, Chong-Gun Kim, and Adam Janiak, editors. Intelligent Information and Database Systems: Third International Conference, ACIIDS 2011, Daegu, Korea, April 20– 22, 2011, Proceedings, Part II, volume 6592 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20041-9 (print), 3-642-20042-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20042-7.

Ng:2006:AKD

[NKLC06] Wee Keong Ng, Masaru Kitsuregawa, Jianzhong Li, and Kuiyu Chang, editors. Advances in Knowledge Discovery and Data Mining: 10th Pacific–Asia Conference, PAKDD 2006, Singapore, April 9–12, 2006. Proceedings, volume 3918 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33206-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33206-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3918.

Nielsen:2009:STP

[NKM^{+}09] Mogens Nielsen, Antonín Kučera, Peter Bro Miltersen, Catuscia Palamidessi, Petr Tůma, and Frank Valencia, editors. SOF-SEM 2009: Theory and Practice of Computer Science: 35th Conference on Current Trends in Theory and Practice of Computer Science, Špindlerův Mlýn, Czech Republic, January 24–30, 2009. Proceedings, volume 5404 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-95890-8 (print), 3-540-95891-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
95891-8.

Ngu:2005:WIS

[NKN^{+}05] Anne H. H. Ngu, Masaru Kitsuregawa, Erich J. Neuhold, Jen-Yao Chung, and Quan Z. Sheng, editors. Web Information Systems Engineering — WISE 2005: 6th International Conference on Web Information Systems Engineering, New York, NY, USA, November 20–22, 2005. Proceedings, volume 3806 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30017-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        30017-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3806.

Nejdl:2008:AHA

[NKPH08] Wolfgang Nejdl, Judy Kay, Pearl Pu, and Eelco Herder, editors. Adaptive Hypermedia and Adaptive Web-Based Systems: 5th International Conference, AH 2008, Hannover, Germany, July 29–August 1, 2008. Proceedings, volume 5149 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CO-DEN LNCSD9. ISBN 3-540-70984-3 (print), 3-540-70987-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I59 I585 2008; QA76 L47 no. 5149. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70984-8;   http://www.
                                        springerlink.com/content/
                                        978-3-540-70987-9.

Nierstrasz:1999:SEE

[NL99] Oscar Marius Nierstrasz and Michel Lemoine, editors. Software engineering— ESEC/FSE ’99: 7th European Software Engineering Conference held jointly with the 7th ACM SIGSOFT Symposium on the Foundations of Software Engineering, Toulouse, France, September 6–10, 1999: Proceedings, volume 1687 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66538-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1687; QA76.758 .E97 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1687.htm;
 http://www.springerlink.
com/content/978-3-540-
66538-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1687.

Nagel:2001:AMD

[NL01] H. H. Nagel and Francisco J. Perales Lopez, editors. Articulated motion and deformable objects: First International Workshop, AMDO 2000, Palma de Mallorca, Spain, September 7–9, 2000: Proceedings, volume 1899 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-67912-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .A486 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1899.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67912-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1899.

Nicholson:2009:AAA

[NL09] Ann Nicholson and Xiaodong Li, editors. AI 2009: Advances in Artificial Intelligence: 22nd Australasian Joint Conference, Melbourne, Australia, December 1–4, 2009. Proceedings, volume 5866 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10438-X (print), 3-642-10439-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10439-8.

Ng:2006:IRT

[NLKJ06] Hwee Tou Ng, Mun-Kew Leong, Min-Yen Kan, and Donghong Ji, editors. Information Retrieval Technology: Third Asia Information Retrieval Symposium, AIRS 2006, Singapore, October 16–18, 2006. Proceedings, volume 4182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45780-1 (print), 3-540-46237-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46237-8.

Nittel:2008:GNS

[NLS08] Silvia Nittel, Alexandros Labrinidis, and Anthony Stefanidis, editors. GeoSensor Networks: Second International Conference, GSN 2006, Boston, MA, USA, October 1–3, 2006, Revised Selected and Invited Papers, volume 4540 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79995-8 (print), 3-540-79996-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79996-2.

Nguyen:2010:IIDa

[NLŚ10a] Ngoc Thanh Nguyen, Manh Thanh Le, and Jerzy Światek, editors. Intelligent Information and Database Systems: Second International Conference, ACI-IDS, Hue City, Vietnam, March 24–26, 2010. Proceedings, Part I, volume 5990 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12144-6 (print), 3-642-12145-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12145-6.

Nguyen:2010:IIDb

[NLŚ10b] Ngoc Thanh Nguyen, Manh Thanh Le, and Jerzy Światek, editors. Intelligent Information and Database Systems: Second International Conference, ACI-IDS, Hue City, Vietnam, March 24–26, 2010. Proceedings, Part II, volume 5991 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12100-4 (print), 3-642-12101-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12101-2.

Nerode:1994:LFC

[NM94] Anil Nerode and Iu. V.

Matiiasevich, editors. Logical foundations of computer science: third International Symposium, LFCS ’94, St. Petersburg, Russia, July 11–14, 1994: proceedings, volume 813 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58140-5 (Berlin), 0-387-58140-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .L64 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0813.htm;
http://www.springerlink.
com/content/978-3-540-
58140-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=813.

Nicholas:1997:IHA

[NM97] Charles K. Nicholas and James Mayfield, editors. Intelligent hypertext: advanced techniques for the World Wide Web, volume 1326 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63637-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 182 pp. LCCN QA76.76.H94I57 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1326.htm;
 http://www.springerlink.
com/content/978-3-540-
63637-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1326.

Nagl:2008:CDC

[NM08] Manfred Nagl and Wolfgang Marquardt, editors. Collaborative and Distributed Chemical Engineering. From Understanding to Substantial Design Process Support: Results of the IMPROVE Project, volume 4970 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70551-1 (print), 3-540-70552-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70552-9.

Nakakoji:2010:NFA

[NMM10] Kumiyo Nakakoji, Yohei Murakami, and Eric Mc-Cready, editors. New Frontiers in Artificial Intelligence: JSAI-isAI 2009 Workshops, LENLS, JURISIN, KCSD, LLLL, Tokyo, Japan, November 19–20, 2009, Revised Selected Papers, volume 6284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14887-5 (print), 3-642-14888-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14888-0.

Nunez:2004:AFM

[NMP^{+}04] Manuel Núñez, Zakaria Maamar, Fernando L. Pelayo, Key Pousttchi, and Fernando Rubio, editors. Applying Formal Methods: Testing, Performance, and M/E-Commerce: FORTE 2004 Workshops: The FormEMC, EPEW, ITM: Toledo, Spain, October 1–2, 2004: Proceedings, volume 3236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23169-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3236.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23169-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3236;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100825.

Nanya:2008:SAI

[NMPM08] Takashi Nanya, Fumihiro Maruyama, András Pataricza, and Miroslaw Malek, editors. Service Availability: 5th International Service Availability Symposium, ISAS 2008 Tokyo, Japan, May 19–21, 2008 Proceedings, volume 5017 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68128-0 (print), 3-540-68129-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
68129-8.

Nitzberg:1993:FSD

[NMS93] M. (Mark) Nitzberg, David Mumford, and Takahiro Shiota. Filtering, segmentation, and depth, volume 662 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56484-5. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 143 pp. LCCN TA1634 .N58 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0662.htm;
http://www.springerlink.
com/content/978-3-540-
56484-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=662.

Nakano:2012:IVA

[NNPW12] Yukiko Nakano, Michael Neff, Ana Paiva, and Marilyn Walker, editors. Intelligent Virtual Agents: 12th International Conference, IVA 2012, Santa Cruz, CA, USA, September, 12–14, 2012. Proceedings, volume 7502 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33196-3 (print), 3-642-33197-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33197-8.

Najm:2003:FMO

[NNS03] Elie Najm, Uwe Nestmann, and Perdita Stevens, editors. Formal Methods for Open Object-Based Distributed Systems: 6th IFIP WG 6.1 International Conference, FMOODS 2003, Paris, France, November 19–21, 2003: Proceedings, volume 2884 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20491-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I573 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2884.
htm;   http://www.springerlink.
com/content/978-3-540-
20491-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2884;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94120.

Narayanan:2006:CVAa

[NNS06a] P. J. Narayanan, Shree K.

Nayar, and Heung-Yeung Shum, editors. Computer Vision — ACCV 2006: 7th Asian Conference on Computer Vision, Hyderabad, India, January 13–16, 2006, Proceedings, Part I, volume 3851 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31219-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31219-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3851.

Narayanan:2006:CVAb

[NNS06b] P. J. Narayanan, Shree K.

Nayar, and Heung-Yeung Shum, editors. Computer Vision — ACCV 2006: 7th Asian Conference on Computer Vision, Hyderabad, India, January 13–16, 2006, Proceedings, Part II, volume 3852 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31244-7 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        31244-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3852.

Natarajan:2011:DCI

[NO11] Raja Natarajan and Adegboyega Ojo, editors. Distributed Computing and Internet Technology: 7th International Conference, ICDCIT 2011, Bhubaneshwar, India, February 9–12, 2011. Proceedings, volume 6536 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19055-3 (print), 3-642-19056-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19056-8.

Noble:2012:EOO

[Nob12] James Noble, editor. ECOOP 2012 — Object-Oriented Programming: 26th European Conference, Beijing, China, June 11–16, 2012. Proceedings, volume 7313 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31056-7 (print), 3-642-31057-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31057-7.

Nokel:1991:TDS

[Nok91] K. (Klaus) Nokel. Temporally distributed symptoms in technical diagnosis, volume 517 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54316-3 (Berlin), 0-387-54316-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 164 pp. LCCN QA76.76.E95 N6 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0517.
htm;   http://www.springerlink.
com/content/978-3-540-
54316-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=517.

Noltemeier:1981:GCC

[Nol81] Hartmut Noltemeier, editor. Graphtheoretic concepts in computer science: proceedings of the International Workshop WG 80, Bad Honnef, June 15–18, 1980, volume 100 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10291-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G74. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0100.htm;
 http://www.springerlink.
com/content/978-0-387-
10291-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=100.

Noltemeier:1988:CGA

[Nol88] Hartmut Noltemeier, editor. Computational geometry and its applications, volume 333 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-50335-8 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA448.D38 I581 1988. DM40.50. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0333.htm;

http://www.springerlink.
com/content/978-0-387-
50335-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=333.

Ng:2005:DIS

[NOOS05] Wee Siong Ng, Beng Chin Ooi, Aris Ouksel, and Claudio Sartori, editors. Databases, information systems, and peer-to-peer computing: second international workshop, DBISP2P 2004, Toronto, Canada, August 29–30, 2004. Revised selected papers, volume 3367 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25233-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.525 .D35 2004. URL http:

//www.springerlink.
com/content/978-3-540-
25233-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3367;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106971.

Nori:1986:FST

[Nor86] Kesav V. Nori, editor.

Foundations of software technology and theoretical computer science: sixth conference, New Delhi, India, December 18–20, 1986: proceedings, volume 241 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-17179-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F681 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0241.htm;
 http://www.springerlink.
com/content/978-0-387-
17179-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=241.

Nori:1987:FST

[Nor87] Kesav V. Nori, editor. Foundations of software technology and theoretical computer science: seventh conference, Pune, India, December 17–19, 1987: proceedings, volume 287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18625-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F681 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0287.htm;
 http://www.springerlink.
com/content/978-0-387-
18625-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=287.

Papers presented at the Seventh Conference on Foundations of Software Technology and Theoretical Computer Science.

Nori:1990:FST

[Nor90] Kesav V. Nori, editor.

Foundations of software technology and theoretical computer science: tenth conference, Bangalore, India, December 17–19, 1990: proceedings, volume 472 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-53487-3 (New York), 3-540-53487-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F68 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0472.htm;
 http://www.springerlink.
com/content/978-0-387-
53487-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=472.

“Tenth Conference on Foundations of Software Technology and Theoretical Computer Science”– Dedication.

North:1997:GDS

[Nor97] Stephen C. North, editor. Graph drawing: Symposium on Graph Drawing, GD ’96, Berkeley, California, USA, September 18–20, 1996: proceedings, volume 1190 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62495-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1190.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62495-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1190.

Nord:2004:SPL

[Nor04] Robert L. Nord, editor. Software Product Lines: Third International Conference, SPLC 2004, Boston, MA, USA, August 30– September 2, 2004. Proceedings, volume 3154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22918-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3154.htm;
 http://www.springerlink.
com/content/978-3-540-
22918-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3154;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100081.

Nossum:1988:ATA

[Nos88] Rolf T. Nossum, editor. Advanced topics in artificial intelligence: 2nd advanced course, ACAI ’87, Oslo, Norway, July 28–August 7, 1987, volume 345 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-50676-4 (USA) (Germany, West). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q335.5 .A38 1988. DM36.00.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t0345.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        50676-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=345.

Organized by European Coordinating Committee for Artificial Intelligence.

Nivat:1985:AIW

[NP85] M. Nivat and Dominique Perrin, editors. Automata on infinite words, volume 192 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15641-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .E261 1984. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0192.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        15641-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=192.

Proceedings of the 1984

spring school.

Naccache:2002:PKC

[NP02] David Naccache and Pascal Paillier, editors. Public key cryptography: 4th [i.e., 5th] International Workshop on Practice and Theory in Public Key Cryptosystems, PKC 2002, Paris, France, February 12–14, 2002: Proceedings, volume 2274 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43168-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I567 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2274.
htm;   http://www.springerlink.
com/content/978-3-540-
43168-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2274.

Nambiar:2009:PEB

[NP09] Raghunath Nambiar and Meikel Poess, editors. Performance Evaluation and Benchmarking: First TPC Technology Conference, TPCTC 2009, Lyon, France, August 24–28, 2009, Revised Selected Papers, volume 5895 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10423-1 (print), 3-642-10424-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10424-4.

Nguyen:2010:PKC

[NP10] Phong Q. Nguyen and David Pointcheval, editors. Public Key Cryptography — PKC 2010: 13th International Conference on Practice and Theory in Public Key Cryptography, Paris, France, May 26– 28, 2010. Proceedings, volume 6056 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13012-7 (print), 3-642-13013-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13013-7.

Nambiar:2011:PEM

[NP11a] Raghunath Nambiar and Meikel Poess, editors. Performance Evaluation, Measurement and Characterization of Complex Systems: Second TPC Technology Conference, TPCTC 2010, Singapore, September 13–17, 2010. Revised Selected Papers, volume 6417 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18205-4 (print), 3-642-18206-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18206-8.

Nitaj:2011:PCA

[NP11b] Abderrahmane Nitaj and David Pointcheval, editors. Progress in Cryptology — AFRICACRYPT 2011: 4th International Conference on Cryptology in Africa, Dakar, Senegal, July 5–7, 2011. Proceedings, volume 6737 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21968-3 (print), 3-642-21969-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21969-6.

Nambiar:2012:TPE

[NP12a] Raghunath Nambiar and Meikel Poess, editors. Topics in Performance Evaluation, Measurement and Characterization: Third TPC Technology Conference, TPCTC 2011, Seattle, WA, USA, August 29– September 3, 2011, Revised Selected Papers, volume 7144 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32626-9 (print), 3-642-32627-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32627-1.

Navarro:2012:SSA

[NP12b] Gonzalo Navarro and Vladimir Pestov, editors. Similarity Search and Applications: 5th International Conference, SISAP 2012, Toronto, ON, Canada, August 9– 10, 2012. Proceedings, volume 7404 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32152-6 (print), 3-642-32153-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
32153-5.

Najm:2006:FTN

[NPPDG06] Elie Najm, Jean-François Pradat-Peyre, and Véronique Viguié Donzeau-Gouge, editors. Formal Techniques for Networked and Distributed Systems

— FORTE 2006: 26th IFIP WG 6.1 International Conference, Paris, France, September 26– 29, 2006. Proceedings, volume 4229 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46219-8 (print), 3-540-46220-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46220-0.

Nishita:2006:ACG

[NPS06] Tomoyuki Nishita, Qunsheng Peng, and Hans-Peter Seidel, editors. Advances in Computer Graphics: 24th Computer Graphics International Conference, CGI 2006, Hangzhou, China, June 26–28, 2006. Proceedings, volume 4035 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35638-X (print), 3-540-35639-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        35639-4.

Nipkow:2002:IHP

[NPW02] Tobias Nipkow, L. C. Paulson, and M. Wenzel. Isabelle/HOL: A Proof Assistant for Higher-Order Logic, volume 2283 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43376-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 218 pp. LCCN QA76.9.L63 I83 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2283.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        43376-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2283.

Ning:2006:ICS

[NQL06] Peng Ning, Sihan Qing, and Ninghui Li, editors. Information and Communications Security: 8th International Conference, ICICS 2006, Raleigh, NC, USA, December 4–7, 2006. Proceedings, volume 4307 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49496-0 (print), 3-540-49497-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49497-3.

Nunez-Queija:2009:NCO

[NQR09] Rudesindo Núñez-Queija and Jacques Resing, editors. Network Control and Optimization: Third Euro-NF Conference, NET-COOP 2009 Eindhoven, The Netherlands, November 23– 25, 2009 Proceedings, volume 5894 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10405-3 (print), 3-642-10406-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10406-0.

Nedellec:1998:MLE

[NR98] Claire Nedellec and Celine Rouveirol, editors. Machine learning: ECML-98: 10th European Conference on Machine Learning, Chemnitz, Germany, April 1998: Proceedings, volume 1398 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64417-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5.E85 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1398.
htm;   http://www.springerlink.
com/content/978-3-540-
64417-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1398.

Narendran:1999:RTA

[NR99] Paliath Narendran and M. Rusinowitch, editors. Rewriting techniques and applications: 10th International Conference, RTA-99, Trento, Italy, July 2–4, 1999: Proceedings, volume 1631 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66201-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .R48 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1631.htm;
 http://www.springerlink.
com/content/978-3-540-
66201-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1631.

Nielsen:2000:MFC

[NR00] Mogens Nielsen and Branislav Rovan, editors. Mathematical Foundations of Computer Science 2000: 25th International Symposium, MFCS 2000, Bratislava, Slovakia, August 28– September 1, 2000: Proceedings, volume 1893 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67901-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1893.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1893.
htm;   http://www.springerlink.
com/content/978-3-540-
67901-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1893.

Nikoletseas:2004:AAW

[NR04] Sotiris Nikoletseas and José D. P. Rolim, editors. Algorithmic Aspects of Wireless Sensor Networks: First International Workshop, Algosensors 2004, Turku, Finland, July 16, 2004, Proceedings, volume 3121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22476-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3121.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22476-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3121;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98740.

Nikoletseas:2006:AAW

[NR06] Sotiris E. Nikoletseas and José D. P. Rolim, editors. Algorithmic Aspects of Wireless Sensor Networks: Second International Workshop, ALGOSENSORS 2006, Venice, Italy, July 15, 2006, Revised Selected Papers, volume 4240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69085-9 (print), 3-540-69087-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69087-0.

Nam:2007:ISC

[NR07] Kil-Hyun Nam and Gwangsoo Rhee, editors. Information Security and Cryptology — ICISC 2007: 10th International Conference, Seoul, Korea, November 29– 30, 2007. Proceedings, volume 4817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76787-8 (print), 3-540-76788-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76788-6.

Nordstrom:2008:ANL

[NR08] Bengt Nordström and Aarne Ranta, editors. Advances in Natural Language Processing: 6th International Conference, GoTAL 2008 Gothenburg, Sweden, August 25–27, 2008 Proceedings, volume 5221 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85286-7 (print), 3-540-85287-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85287-2.

Nijholt:2012:ACE

[NRR12] Anton Nijholt, Teresa Romão, and Dennis Reidsma, editors. Advances in Computer Entertainment: 9th International Conference, ACE 2012, Kathmandu, Nepal, November 3–5, 2012. Proceedings, volume 7624 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34291-4 (print), 3-642-34292-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34292-9.

Nardi:2005:RRS

[NRSSV05] Daniele Nardi, Martin Riedmiller, Claude Sammut, and José Santos-Victor, editors. RoboCup 2004: Robot Soccer World Cup VIII: Robot Soccer World Cup VIII, volume 3276 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25046-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25046-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3276.

Nickles:2004:ACA

[NRW04] Matthias Nickles, Michael Rovatsos, and Gerhard Weiss, editors. Agents and Computational Autonomy: Potential, Risks, and Solutions, volume 2969 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22477-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 A445 2004.

URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t2969.
htm;   http://www.springerlink.
com/content/978-3-540-
22477-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2969;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/

b99010. This volume contains the postproceedings of the 1st International Workshop on Computational Autonomy Potential, Risks, Solutions (AUTONOMY 2003), held at the 2nd International Joint Conference on Autonomous Agents and Multi-agent Systems (AAMAS 2003), July 14, 2003, Melbourne, Australia.

Nieto:2009:ATC

[NRWI09] Juan González Nieto, Wolfgang Reif, Guojun Wang, and Jadwiga Indulska, editors. Autonomic and Trusted Computing: 6th International Conference, ATC 2009 Brisbane, Australia, July 7–9, 2009 Proceedings, volume 5586 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02703-2 (print), 3-642-02704-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02704-8.

Nishizeki:1981:GTA

[NS81] T. Nishizeki and N. Saito, editors. Graph theory and algorithms: 17th symposium of Research Institute of Electrical Communication, Tohoku University, Sendai, Japan, October 24– 25, 1980: proceedings, volume 108 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10704-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G72. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0108.htm;
http://www.springerlink.
com/content/978-0-387-
10704-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=108.

Nielsen:1982:ALP

[NS82] M. Nielsen and E. M.

Schmidt, editors. Automata, languages, and programming: ninth colloquium, Aarhus, Denmark, July 12–16, 1982, volume 140 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11576-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.140. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0140.htm;
 http://www.springerlink.
com/content/978-0-387-
11576-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=140.

Nichols:1987:EES

[NS87] H. K. Nichols and D. Simpson, editors. ESEC ’87: 1st European Software Engineering Conference, Strasbourg, France, September 9–11, 1987: proceedings, volume 289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18712-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .E971 1987. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t0289.htm;
http://www.springerlink.
com/content/978-0-387-
18712-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=289.

Norrie:1990:CAL

[NS90] D. H. Norrie and H.-W.

Six, editors. Computer assisted learning: 3rd International Conference, ICCAL ’90, Hagen, FRG, June 11–13, 1990 proceedings, volume 438 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52699-4 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.5 .I1163 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0438.htm;
http://www.springerlink.
com/content/978-0-387-
52699-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=438.

Nikolaou:1998:RAT

[NS98] Christos Nikolaou and Constantine Stephanidis, editors. Research and Advanced Technology for Digital Libraries: Second European Conference, ECDL ’98, Heraklion, Crete, Cyprus, September 21–23, 1998: Proceedings, volume 1513 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65101-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4082.E85E25 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1513.htm;
 http://www.springerlink.
com/content/978-3-540-
65101-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1513.

Noriega:1999:AME

[NS99] Pablo Noriega and Carles Sierra, editors. Agent mediated electronic commerce: First International Workshop on Agent-Mediated Electronic Trading, AMET-98: Minneapolis, MN, USA, May 10th, 1998: Selected Papers, volume 1571 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65955-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN HF5548.32 .I585 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1571.htm;
http://www.springerlink.
com/content/978-3-540-
65955-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1571.

Nielsen:2000:ATP

[NS00] M. Nielsen and D. Simpson, editors. Application and theory of Petri nets 2000: 21st International Conference, ICATPN 2000, Aarhus, Denmark, June 26–30, 2000: Proceedings, volume 1825 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67693-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1825.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1825.
htm;   http://www.springerlink.
com/content/978-3-540-
67693-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1825.

Nordahl:2010:HAI

[NSFB10] Rolf Nordahl, Stefania Serafin, Federico Fontana, and Stephen Brewster, editors. Haptic and Audio Interaction Design: 5th International Workshop, HAID 2010, Copenhagen, Denmark, September 16– 17, 2010. Proceedings, volume 6306 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15840-4 (print), 3-642-15841-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15841-4.

Nunzi:2009:IOM

[NSL09] Giorgio Nunzi, Caterina Scoglio, and Xing Li, editors. IP Operations and Management: 9th IEEE International Workshop, IPOM 2009, Venice, Italy, October 29–30, 2009. Proceedings, volume 5843 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04967-2 (print), 3-642-04968-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
04968-2.

Nagl:2000:AGT

[NSM00] Manfred Nagl, Andreas Schürr, and Manfred Munch, editors. Applications of graph transformations with industrial relevance: International Workshop, AG-TIVE’99, Kerkrade, The Netherlands, September 1–3, 1999: Proceedings, volume 1779 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67658-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .A195 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1779.htm;
http://www.springerlink.
com/content/978-3-540-
67658-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1779.

Neves:2007:PAI

[NSM07] José Neves, Manuel Filipe Santos, and José Manuel Machado, editors. Progress in Artificial Intelligence: 13th Portuguese Conference on Artificial Intelligence, EPIA 2007, Workshops: GAIW, AIASTS, ALEA, AMITA, BAOSW, BI, CMBSB, IROBOT, MASTA, STCS, and TEMA, Guimarães, Portugal, December 3– 7, 2007. Proceedings, volume 4874 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77000-3 (print), 3-540-77002-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77002-2.

Nystrom:2003:DGC

[NSS03] Ingela Nyström, Gabriella Sanniti di Baja, and Stina Svensson, editors. Discrete Geometry for Computer Imagery: 11th International Conference, DGCI 2003, Naples, Italy, November 19–21, 2003: Proceedings, volume 2886 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20499-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .D525 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2886.htm;

http://www.springerlink.
com/content/978-3-540-
20499-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2886;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94107.

Nasraoui:2007:AWM

[NSS^{+}07] Olfa Nasraoui, Myra Spiliopoulou, Jaideep Srivastava, Bamshad Mobasher, and Brij Masand, editors. Advances in Web Mining and Web Usage Analysis: 8th International Workshop on Knowledge Discovery on the Web, WebKDD 2006 Philadelphia, USA, August 20, 2006 Revised Papers, volume 4811 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77484-X (print), 3-540-77485-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77485-3.

Nerode:1992:LFC

[NT92] Anil Nerode and M. A.

Taitslin, editors. Logical foundations of computer science — Tver ’92: Second International Symposium, Tver, Russia, July 20–24, 1992 proceedings, volume 620 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55707-5 (Berlin), 0-387-55707-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .L64 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0620.htm;
 http://www.springerlink.
com/content/978-3-540-
55707-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=620.

Nielsen:1998:CSL

[NT98] Mogens Nielsen and Wolfgang Thomas, editors. Computer science logic: 11th International Workshop, CSL ’97: Annual Conference of the EACSL, Aarhus, Denmark, August 23– 29, 1997: Selected Papers, volume 1414 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64570-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1414. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1414.htm;
 http://www.springerlink.
com/content/978-3-540-
64570-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1414.

Nixon:2003:TMF

[NT03] Paddy Nixon and Sotirios Terzis, editors. Trust Management: First International Conference, iTrust 2003, Heraklion, Crete, Greece, May 28– 30, 2003: Proceedings, volume 2692 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40224-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 I38 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2692.htm;
 http://www.springerlink.
com/content/978-3-540-
40224-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2692.

Also available via the World Wide Web.

Nejdl:2006:IAL

[NT06] Wolfgang Nejdl and Klaus Tochtermann, editors. Innovative Approaches for Learning and Knowledge Sharing: First European Conference on Technology Enhanced Learning, EC-TEL 2006 Crete, Greece, October 1–4, 2006 Proceedings, volume 4227 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45777-1 (print), 3-540-46234-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46234-7.

Nurmi:1992:ATS

[NU92] O. Nurmi and E. Ukkonen, editors. Algorithm theory: SWAT ’92: Third Scandinavian Workshop on Algorithm Theory, Helsinki, Finland, July 8–10, 1992: proceedings, volume 621 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55706-7 (Berlin), 0-387-55706-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .S28 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0621.htm;
 http://www.springerlink.
com/content/978-3-540-
55706-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=621.

Nurnberg:2003:MIS

[Nür03] Peter J. Nürnberg, editor. Metainformatics: International Symposium, MIS 2002, Esbjerg, Denmark, August 7–10, 2002: Revised Papers, volume 2641 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40218-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .M55 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2641.htm;
 http://www.springerlink.
com/content/978-3-540-
40218-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2641.

Also available via the World Wide Web.

Niessen:2001:MIC

[NV01a] Wiro J. Niessen and M. A. Viergever, editors. Medical image computing and computerassisted intervention-MICCAI 2001: 4th International Conference, Utrecht, The Netherlands, October 14– 17, 2001: Proceedings, volume 2208 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42697-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R857.O6 I574 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2208.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42697-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2208.

Nieuwenhuis:2001:LPA

[NV01b] Robert Nieuwenhuis and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 8th International Conference, LPAR 2001, Havana, Cuba, December 3–7, 2001. Proceedings, volume 2250 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42957-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I55 2001; QA267.A1 L43 no.2250. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2250.htm;
 http://www.springerlink.
com/content/978-3-540-
42957-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2250.

Nelson:2008:CGI

[NV08] Craig E. Nelson and Stéphane Vialette, editors. Comparative Genomics: International Workshop, RECOMB-CG 2008, Paris, France, October 13–15, 2008. Proceedings, volume 5267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87988-9 (print), 3-540-87989-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87989-3.

Nouwen:2011:VCI

[NvRSS11] Rick Nouwen, Robert van Rooij, Uli Sauerland, and Hans-Christian Schmitz, editors. Vagueness in Communication: International Workshop, ViC 2009, held as part of ESSLLI 2009, Bordeaux, France, July 20–24, 2009. Revised Selected Papers, volume 6517 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18445-6 (print), 3-642-18446-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18446-8.

Noriega:2007:COI

[NVSB^{+}07] Pablo Noriega, Javier Vázquez-Salceda, Guido Boella, Olivier Boissier, Virginia Dignum, Nicoletta Fornara, and Eric Matson, editors. Coordination, Organizations, Institutions, and Norms in Agent Systems II: AAMAS 2006 and ECAI 2006 International Workshops, COIN 2006 Hakodate, Japan, May 9, 2006 Riva del Garda, Italy, August 28, 2006. Revised Selected Papers, volume 4386 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74457-6 (print), 3-540-74459-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74459-7.

Nicholas:1997:PDP

[NW97] Charles K. Nicholas and Derick Wood, editors. Principles of document processing: Third International Workshop, PODP’96, Palo Alto, California, USA, September 23, 1996: proceedings, volume 1293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63620-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .P63 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1293.htm;
 http://www.springerlink.
com/content/978-3-540-
63620-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1293.

Naher:2001:AEI

[NW01] Stefan Näher and Dorothea Wagner, editors. Algorithm engineering: 4th International Workshop, WAE 2000, Saarbrucken, Germany, September 5– 8, 2000: Proceedings, volume 1982 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42512-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I59 2000; QA267.A1 L43 no.1982.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1982.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        42512-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1982.

Nikolaidis:2010:AHM

[NW10] Ioanis Nikolaidis and Kui Wu, editors. Ad-Hoc, Mobile and Wireless Networks: 9th International Conference, ADHOC-NOW 2010, Edmonton, AB, Canada, August 20– 22, 2010. Proceedings, volume 6288 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14784-4 (print), 3-642-14785-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14785-2.

Nierstrasz:2006:MDE

[NWHR06] Oscar Nierstrasz, Jon Whittle, David Harel, and Gianna Reggio, editors. Model Driven Engineering Languages and Systems: 9th International Conference, MoD-ELS 2006, Genova, Italy, October 1–6, 2006. Proceedings, volume 4199 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45772-0 (print), 3-540-45773-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45773-2.

Nagel:2006:EPP

[NWL06] Wolfgang E. Nagel, Wolfgang V. Walter, and Wolfgang Lehner, editors. Euro-Par 2006 Parallel Processing: 12th International Euro-Par Conference, Dresden, Germany, August 28 – September 1, 2006. Proceedings, volume 4128 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37783-2 (print), 3-540-37784-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37784-9.

Nakajima:1983:IPS

[NY83] R. Nakajima and T. Yuasa. The IOTA programming system: a modular programming environment, volume 160 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CODEN LNCSD9. ISBN 0-387-12693-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 217 pp. LCCN QA76.73 .I6 1983. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0160.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        12693-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=160.

Nishio:1993:OTA

[NY93] Shojiro Nishio and Akinori Yonezawa, editors. Object technologies for advanced software: first JSSST International Symposium, Kanazawa, Japan, November 4– 6, 1993: proceedings, volume 742 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-57342-9 (USA), 3-540-57342-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .I55 1993. DM104.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0742.
htm;   http://www.springerlink.
com/content/978-0-387-
57342-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=742.

Nyberg:1998:ACE

[Nyb98] Kaisa Nyberg, editor.

Advances in cryptology: EUROCRYPT ’98: International Conference on the theory and application of cryptographic techniques, Espoo, Finland, May 31 — June 4, 1998: Proceedings, volume 1403 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64518-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1403. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1403.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64518-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1403.

Nyberg:2008:FSE

[Nyb08] Kaisa Nyberg, editor. Fast Software Encryption: 15th International Workshop, FSE 2008, Lausanne, Switzerland, February 10–13, 2008, Revised Selected Papers, volume 5086 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-71038-8 (print), 3-540-71039-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71039-4.

Namjoshi:2007:ATV

[NYHO07] Kedar S. Namjoshi, Tomohiro Yoneda, Teruo Higashino, and Yoshio Okamura, editors. Automated Technology for Verification and Analysis: 5th International Symposium, ATVA 2007 Tokyo, Japan, October 22–25, 2007 Proceedings, volume 4762 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75595-0 (print), 3-540-75596-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75596-8.

Nielsen:2003:SQA

[NZ03] Thomas Dyhre Nielsen and Nevin Lianwen Zhang, editors. Symbolic and quantitative approaches to reasoning with uncertainty: 7th European Conference, ECSQARU 2003, Aalborg, Denmark, July 2–5, 2003: Proceedings, volume 2711 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40494-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q375 .E87 2003.

URL http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t2711.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40494-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2711.

Nayak:2006:KDX

[NZ06] Richi Nayak and Mohammed J. Zaki, editors. Knowledge Discovery from XML Documents: First International Workshop, KDXD 2006, Singapore, April 9, 2006. Proceedings, volume 3915 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33180-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33180-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3915.

Nasraoui:2006:AWM

[NZS^{+}06] Olfa Nasraoui, Osmar Zaïane, Myra Spiliopoulou, Bamshad Mobasher, Brij Masand, and Philip S. Yu, editors. Advances in Web Mining and Web Usage Analysis: 7th International Workshop on Knowledge Discovery on the Web, WebKDD 2005, Chicago, IL, USA, August 21, 2005. Revised Papers, volume 4198 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46346-1 (print), 3-540-46348-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46348-1.

Namjoshi:2011:HSV

[NZZ11] Kedar Namjoshi, Andreas Zeller, and Avi Ziv, editors. Hardware and Software: Verification and Testing: 5th International Haifa Verification Conference, HVC 2009, Haifa, Israel, October 19–22, 2009, Revised Selected Papers, volume 6405 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19236-X (print), 3-642-19237-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19237-1.

Ooi:1993:ASD

[OA93] Beng Chin Ooi and David Abel, editors. Advances in spatial databases: Third International Symposium, SSD ’93, Singapore, June 23–25, 1993: proceedings, volume 692 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56869-7 (Berlin), 0-387-56869-7 (New Uork). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S978 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0692.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        56869-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=692.

Okuno:2007:NTA

[OA07] Hiroshi G. Okuno and Moonis Ali, editors. New Trends in Applied Artificial Intelligence: 20th International Conference on Industrial, Engineering and Other Applications of Applied Intelligent Systems, IEA/ AIE 2007, Kyoto, Japan, June 26–29, 2007. Proceedings, volume 4570 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73322-1 (print), 3-540-73325-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73325-6.

Ojeda-Aciego:2000:LAI

[OAdGBP00] Manuel Ojeda-Aciego, Inma P. de Guzmán, Gerhard Brewka, and Luís Moniz Pereira, editors. Logics in artificial intelligence: European workshop, JELIA 2000, Málaga, Spain, September 29–October 2, 2000: Proceedings, volume 1919 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41131-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1919. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1919.htm;
 http://www.springerlink.
com/content/978-3-540-
41131-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1919.

Ortiz-Arroyo:2008:ISI

[OALZ^{+}08] Daniel Ortiz-Arroyo, Henrik Legind Larsen, Daniel Dajun Zeng, David Hicks, and Gerhard Wagner, editors. Intelligence and Security Informatics: First European Conference, EuroISI 2008, Esbjerg, Denmark, December 3–5, 2008. Proceedings, volume 5376 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89899-9 (print), 3-540-89900-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89900-6.

Oakley:2007:HAI

[OB07] Ian Oakley and Stephen Brewster, editors. Haptic and Audio Interaction Design: Second International Workshop, HAID 2007 Seoul, South Korea, November 29–30, 2007 Proceedings, volume 4813 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76701-0 (print), 3-540-76702-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76702-2.

OBoyle:2012:CCI

[O’B12] Michael O’Boyle, editor.

Compiler Construction: 21st International Conference, CC 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7210 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28651-8 (print), 3-642-28652-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28652-0.

Onada:2011:NFA

[OBM11] Takashi Onada, Daisuke Bekki, and Eric Mc-Cready, editors. New Frontiers in Artificial Intelligence: JSAI-isAI 2010 Workshops, LENLS, JURISIN, AMBN, ISS, Tokyo, Japan, November 18–19, 2010, Revised Selected Papers, volume 6797 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25654-6 (print), 3-642-25655-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25655-4.

Okumura:2012:NFA

[OBS12] Manabu Okumura, Daisuke Bekki, and Ken Satoh, editors. New Frontiers in Artificial Intelligence: JSAI-isAI 2011 Workshops, LENLS, JU-RISIN, ALSIP, MiMI, Takamatsu, Japan, December 1–2, 2011. Revised Selected Papers, volume 7258 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32089-9 (print), 3-642-32090-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32090-3.

ODonnell:1977:CSD

[O’D77] Michael J. O’Donnell.

Computing in systems described by equations, volume 58 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CODEN LNCSD9. ISBN 0-387-08531-9. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 111 pp. LCCN QA76.6 .O36. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0058.htm;
 http://www.springerlink.
com/content/978-0-387-
08531-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=58.

Programming languages (Electronic computers).

Ortmeier:2012:CSRa

[OD12a] Frank Ortmeier and Peter Daniel, editors. Computer Safety, Reliability, and Security: 31st International Conference, SAFECOMP 2012, Magdeburg, Germany, September 25–28, 2012. Proceedings, volume 7612 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33677-9 (print), 3-642-33678-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33678-2.

Ortmeier:2012:CSRb

[OD12b] Frank Ortmeier and Peter Daniel, editors. Computer Safety, Reliability, and Security: SAFE-COMP 2012 Workshops: Sassur, ASCoMS, DESEC4LCCI, ERCIM/ EWICS, IWDE, Magdeburg, Germany, September 25–28, 2012. Proceedings, volume 7613 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33674-4 (print), 3-642-33675-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33675-1.

Odersky:2004:EOO

[Ode04] Martin Odersky, editor. ECOOP 2004 — Object-Oriented Programming: 8th European Conference, Oslo, Norway, June 14–18, 2004: Proceedings, volume 3086 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22159-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.O35E33 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3086.
htm;   http://www.springerlink.
com/content/978-3-540-
22159-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3086;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98195.

Odlyzko:1987:ACC

[Odl87] Andrew Michael Odlyzko, editor. Advances in cryptology: CRYPTO ’86: proceedings, volume 263 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 3-540-18047-8, 0-387-18047-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C791 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0263.htm;
 http://www.springerlink.
com/content/978-3-540-
18047-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=263.

Conference held at the University of California, Santa Barbara, Aug. 11– 15, 1986.

Obayashi:2007:EMC

[ODP^{+}07] Shigeru Obayashi, Kalyanmoy Deb, Carlo Poloni, Tomoyuki Hiroyasu, and Tadahiko Murata, editors. Evolutionary Multi-Criterion Optimization: 4th International Conference, EMO 2007, Matsushima, Japan, March 5–8, 2007. Proceedings, volume 4403 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70927-4 (print), 3-540-70928-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70928-2.

Ong:2012:LLI

[OdQ12] Luke Ong and Ruy de Queiroz, editors. Logic, Language, Information and Computation: 19th International Workshop, WoL-LIC 2012, Buenos Aires, Argentina, September 3– 6, 2012. Proceedings, volume 7456 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32620-X (print), 3-642-32621-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32621-9.

Ostrovsky:2008:SCN

[ODV08] Rafail Ostrovsky, Roberto De Prisco, and Ivan Visconti, editors. Security and Cryptography for Networks: 6th International Conference, SCN 2008, Amalfi, Italy, September 10–12, 2008. Proceedings, volume 5229 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85854-7 (print), 3-540-85855-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85855-3.

Oswald:2015:ACE

[OF15] Elisabeth Oswald and Marc Fischlin, editors. Advances in Cryptology — EUROCRYPT 2015: 34th Annual International Conference on the Theory and Applications of Cryptographic Techniques, Sofia, Bulgaria, April 26–30, 2015, Proceedings, Part II, volume 9057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2015. ISBN 3-662-46802-6 (print), 3-662-46803-4 (e-book). LCCN QA76.9.A25. URL http:

//www.springerlink.
com/content/978-3-662-
46803-6.

Olsen:2005:DSS

[OFK05] Ole Fogh Olsen, Luc M. J. Florack, and Arjan Kuijper, editors. Deep structure, singularities, and computer vision: first international workshop, DSSCV 2005, Maastricht, The Netherlands, June 9–10, 2005: revised selected papers, volume 3753 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29836-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .I57 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29836-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3753.

Oddi:2009:RAC

[OFR09] Angelo Oddi, François Fages, and Francesca Rossi, editors. Recent Advances in Constraints: 13th Annual ERCIM International Workshop on Constraint Solving and Constraint Logic Programming, CSCLP 2008, Rome, Italy, June 18–20, 2008, Revised Selected Papers, volume 5655 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03250-8 (print), 3-642-03251-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03251-6.

Odell:2005:AOS

[OGM05] James Odell, Paolo Giorgini, and Jörg P. Müller, editors. Agentoriented software engineering V: 5th international workshop, AOSE 2004, New York, NY, USA, July 19, 2004. Revised selected papers, volume 3382 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24286-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758. URL

                                        http://www.springerlink.
                                        com/content/978-3-540-
                                        24286-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3382;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b105022.

OHallaron:1998:LCR

[O’H98] David O’Hallaron, editor. Languages, Compilers, and Run-Time Systems for Scalable Computers: 4th International Workshop, LCR’98, Pittsburgh, PA, USA, May 28–30, 1998. Selected Papers, volume 1511 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65172-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W658 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1511.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65172-7;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1511.

Ohlbach:1993:GAA

[Ohl93] Hans Jurgen Ohlbach, editor. GWAI-92: advances in artificial intelligence: 16th German Conference on Artificial Intelligence, Bonn, Germany, August 31– September 3, 1992: proceedings, volume 671 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56667-8 (New York), 3-540-56667-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G466 1992. DM80.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0671.htm;
http://www.springerlink.
com/content/978-0-387-
56667-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=671.

Ohori:2003:PLS

[Oho03] Atsushi Ohori, editor. Programming Languages and Systems: First Asian Symposium, APLAS 2003, Beijing, China, November 27– 29, 2003: Proceedings, volume 2895 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20536-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2895.htm;
 http://www.springerlink.
com/content/978-3-540-
20536-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2895;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94290.

Ohta:1998:ACA

[OiDP98] Kazuo Ohta and Ting i (Dingyi) Pei, editors. Advances in Cryptology—ASIACRYPT’98: International Conference on the Theory and Application of Cryptology and Information Security, Beijing, China, October 118-22, 998. Proceedings, volume 1514 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65109-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.9.A25I5553 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1514.htm;
http://www.springerlink.
com/content/978-3-540-
65109-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1514.

Onodera:1990:FMV

[OK90] Tamiya Onodera and Satoru Kawai. A formal model of visualization in computer graphics systems, volume 421 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52395-2 (Berlin), 0-387-52395-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 100 pp. LCCN T385 .O56 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0421.htm;
http://www.springerlink.
com/content/978-3-540-
52395-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=421.

Oren:1996:CAS

[OK96] Tuncer I. Oren and George J. Klir, editors. Computer aided systems theory: CAST’94, 4th International Workshop, Ottawa, Ontario, Canada, May 16–20, 1994: Selected Papers, volume 1105 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61478-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .C642414 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1105.htm;
 http://www.springerlink.
com/content/978-3-540-
61478-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1105.

Okamoto:2000:ACA

[Oka00] Tatsuaki Okamoto, editor. Advances in Cryptology—ASIACRYPT 2000: 6th International Conference on the Theory and Application of Cryptology and Information Security, Kyoto, Japan, December 3–7, 2000. Proceedings, volume 1976 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41404-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I555 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1976.htm;
http://www.springerlink.
com/content/978-3-540-
41404-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1976.

Okada:2003:SST

[Oka03] M. Okada, editor. Software security — theories and systems: Mext-NSF-JSPS International Symposium, ISSS 2002, Tokyo, Japan, November 8–10, 2002: Revised Papers, volume 2609 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00708-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I86 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2609.htm;
http://www.springerlink.
com/content/978-3-540-
00708-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2609.

Okamoto:2004:TCC

[Oka04] Tatsuaki Okamoto, editor. Topics in Cryptology— CT-RSA 2004: The Cryptographers’ Track at the RSA Conference 2004, San Francisco, CA, USA, February 23–27, 2004: Proceedings, volume 2964 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20996-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C822 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2964.
htm;   http://www.springerlink.
com/content/978-3-540-
20996-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2964;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95630.

Osipov:2010:WWI

[OKBMB10] Evgeny Osipov, Andreas Kassler, Thomas Michael Bohnert, and Xavier Masip-Bruin, editors. Wired/Wireless Internet Communications: 8th International Conference, WWIC 2010, Luleå, Sweden, June 1– 3, 2010. Proceedings, volume 6074 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13314-2 (print), 3-642-13315-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13315-2.

Ono:2009:LLI

[OKdQ09] Hiroakira Ono, Makoto Kanazawa, and Ruy de Queiroz, editors. Logic, Language, Information and Computation: 16th International Workshop, WoL-LIC 2009, Tokyo, Japan, June 21–24, 2009. Proceedings, volume 5514 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02260-X (print), 3-642-02261-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02261-6.

Oinas-Kukkonen:2008:PTT

[OKHH^{+}08] Harri Oinas-Kukkonen, Per Hasle, Marja Harjumaa, Katarina Segerståhl, and Peter Øhrstrøm, editors. Persuasive Technology: Third International Conference, PER-SUASIVE 2008, Oulu, Finland, June 4–6, 2008. Proceedings, volume 5033 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68500-6 (print), 3-540-68504-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68504-3.

Owe:2004:OOF

[OKL04] Olaf Owe, Stein Krogdahl, and Tom Lyche, editors. From Object-Orientation to Formal Methods: Essays in Memory of Ole-Johan Dahl, volume 2635 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21366-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2635.htm;
 http://www.springerlink.
com/content/978-3-540-
21366-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

0302-9743&volume=2635;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96089.

Oivo:2002:PFS

[OKS02] Markku Oivo and Seija Komi-Sirviö, editors. Product Focused Software Process Improvement: 14th International Conference, PRO-FES 2002 Rovaniemi, Finland, December 9– 11, 2002. Proceedings, volume 2559 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00234-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I5743 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2559.htm;
http://www.springerlink.
com/content/978-3-540-
00234-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2559.

Oliveira:2000:GIA

[Oli00] Arlindo L. Oliveira, editor. Grammatical inference: algorithms and applications: 5th international colloquium, ICGI 2000, Lisbon, Portugal, September 11–13, 2000: Proceedings, volume 1891 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41011-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1891. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1891.htm;
 http://www.springerlink.
com/content/978-3-540-
41011-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1891;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b75249.

Olivetti:2007:ARA

[Oli07] Nicola Olivetti, editor. Automated Reasoning with Analytic Tableaux and Related Methods: 16th International Conference, TABLEAUX 2007, Aix en Provence, France, July 3–6, 2007. Proceedings, volume 4548 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73098-2 (print), 3-540-73099-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73099-6.

Ogier:2010:GRA

[OLL10] Jean-Marc Ogier, Wenyin Liu, and Josep Lladós, editors. Graphics Recognition. Achievements, Challenges, and Evolution: 8th International Workshop, GREC 2009, La Rochelle, France, July 22–23, 2009. Selected Papers, volume 6020 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13727-X (print), 3-642-13728-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13728-0.

Olthoff:1995:EOO

[Olt95] Walter Olthoff, editor.

ECOOP ’95, objectoriented programming: 9th European Conference, Aarhus, Denmark, August 7–11, 1995: proceedings, volume 952 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60160-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0952.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60160-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=952.

Olveczky:2010:RLA

[Ölv10] Peter Csaba Ölveczky, editor. Rewriting Logic and Its Applications: 8th International Workshop, WRLA 2010, Held as a Satellite Event of ETAPS 2010, Paphos, Cyprus, March 20–21, 2010, Revised Selected Papers, volume 6381 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16309-2 (print), 3-642-16310-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65. URL

                                        http://www.springerlink.
                                        com/content/978-3-642-
                                        16310-4.

Orr:1998:NNT

[OM98] Genevieve B. Orr and Klaus-Robert Müller, editors. Neural Networks: Tricks of the Trade, volume 1524 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65311-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 432 pp. LCCN QA76.87.N4913 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1524.htm;
 http://www.springerlink.
com/content/978-3-540-
65311-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1524.

Omidyar:2000:MWC

[Omi00] Cambyse Guy Omidyar, editor. Mobile and wireless communications networks: IFIP-TC6/European Commission Networking 2000 International Workshop, MWCN 2000: Paris, France, May 16–17, 2000: Proceedings, volume 1818 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67543-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN TK5103.2 .I347 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1818.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67543-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1818.

Oliveira:2006:BMD

[OMMSP06] José Luís Oliveira, Víctor Maojo, Fernando Martin-Sanchez, and António Sousa Pereira, editors. Biological and Medical Data Analysis: 6th International Symposium, IS-BMDA 2005, Aveiro, Portugal, November 10– 11, 2005. Proceedings, volume 3745 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-29674-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29674-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3745.

Osaki:1980:RES

[ON80] Shunji Osaki and Toshihiko Nishio. Reliability evaluation of some faulttolerant computer architectures, volume 97 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10274-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 129 pp. LCCN QA76.9.F38 .O78. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0097.htm;
 http://www.springerlink.
com/content/978-0-387-
10274-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=97.

Errata slip inserted.

OShea:2011:AMA

[ONC^{+}11] James O’Shea, Ngoc Thanh Nguyen, Keeley Crockett, Robert J. Howlett, and Lakhmi C. Jain, editors. Agent and Multi-Agent Systems: Technologies and Applications: 5th KES International Conference, KES-AMSTA 2011, Manchester, UK, June 29 — July 1, 2011. Proceedings, volume 6682 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21999-3 (print), 3-642-22000-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22000-5.

Ong:2005:CSL

[Ong05] Luke Ong, editor. Computer Science Logic: 19th International Workshop, CSL 2005, 14th Annual Conference of the EACSL, Oxford, UK, August 22–25, 2005: Proceedings, volume 3634 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28231-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.L63 W67 2005. URL http://

                                        springerlink.metapress.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3634;   http:
                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28231-0.

Ong:2010:FSS

[Ong10] Luke Ong, editor. Foundations of Software Science and Computational Structures: 13th International Conference, FOS-SACS 2010, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2010, Paphos, Cyprus, March 20–28, 2010. Proceedings, volume 6014 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12031-8 (print), 3-642-12032-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12032-9.

Ong:2011:TLC

[Ong11] Luke Ong, editor. Typed Lambda Calculi and Applications: 10th International Conference, TLCA 2011, Novi Sad, Serbia, June 1–3, 2011. Proceedings, volume 6690 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21690-0 (print), 3-642-21691-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21691-6.

Obermaisser:2007:STE

[ONPR07]  Roman  Obermaisser,

Yunmook Nah, Peter Puschner, and Franz J. Rammig, editors. Software Technologies for Embedded and Ubiquitous Systems: 5th IFIP WG 10.2 International Workshop, SEUS 2007, Santorini Island, Greece, May 2007. Revised Papers, volume 4761 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75663-9 (print), 3-540-75664-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75664-4.

Ober:2012:SIS

[OO12] Iulian Ober and Ileana Ober, editors. SDL 2011: Integrating System and Software Modeling: 15th International SDL Forum Toulouse, France, July 5–7, 2011. Revised Papers, volume 7083 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25263-X (print), 3-642-25264-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25264-8.

Ooi:1990:EQP

[Ooi90] Beng Chin Ooi. Efficient query processing in geographic information systems, volume 471 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-53474-1 (Berlin), 0-387-53474-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 208 pp. LCCN G70.2 .O55 1990. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0471.
htm;   http://www.springerlink.
com/content/978-3-540-
53474-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=471.

Based on the author’s Ph.D. dissertation.

Olive:1997:AIS

[OP97] A. Olive and Joan Antoni Pastor, editors. Advanced information systems engineering: 9th International Conference, CAiSE’97, Barcelona, Catalonia, Spain, June 16–20, 1997: proceedings, volume 1250 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63107-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I57115 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1250.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63107-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1250.

Oza:2005:MCS

[OPKR05] Nikunj C. Oza, Robi Polikar, Josef Kittler, and Fabio Roli, editors. Multiple classifier systems: 6th international workshop, MCS 2005, Seaside, CA, USA, June 13–15, 2005, proceedings, volume 3541 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26306-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26306-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3541.

Omicini:2004:ESA

[OPP04] Andrea Omicini, Paolo Petta, and Jeremy Pitt, editors. Engineering Societies in the Agents World: 4th International Workshop, ESAW 2003, London, UK, October 29–31, 2003: Revised Selected and Invited Papers, volume 3071 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22231-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 E83 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3071.
htm;   http://www.springerlink.
com/content/978-3-540-
22231-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3071;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98212.

Olken:2011:RBM

[OPS11] Frank Olken, Monica Palmirani, and Davide Sottara, editors. Rule

— Based Modeling and Computing on the Semantic Web: 5th International Symposium, RuleML 2011– America, Ft. Lauderdale, FL, Florida, USA, November 3–5, 2011. Proceedings, volume 7018 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24907-8 (print), 3-642-24908-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24908-2.

Omicini:2001:ESAb

[OPT01] Andrea Omicini, Paolo Petta, and Robert Tolksdorf, editors. Engineering Societies in the Agents World II: Second International Workshop, ESAW 2001, Prague, Czech Republic, July 7, 2001: Revised Papers, volume 2203 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-43091-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 E83 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2203.
htm;   http://www.springerlink.

com/content/978-3-540-
43091-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2203.

Oquendo:2003:SPT

[Oqu03] Flavio Oquendo, editor. Software Process Technology: 9th European Workshop, EWSPT 2003, Helsinki, Finland, September 1–2, 2003: Proceedings, volume 2786 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40764-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 E87 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2786.
htm;   http://www.springerlink.
com/content/978-3-540-
40764-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2786;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12019.

Oquendo:2007:SAF

[Oqu07] Flavio Oquendo, editor.

Software Architecture: First European Conference, ECSA 2007 Aranjuez, Spain, September 24–26, 2007 Proceedings, volume 4758 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75131-9 (print), 3-540-75132-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75132-8.

Oswald:2008:CHE

[OR08] Elisabeth Oswald and Pankaj Rohatgi, editors. Cryptographic Hardware and Embedded Systems — CHES 2008: 10th International Workshop, Washington, D.C., USA, August 10–13, 2008. Proceedings, volume 5154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85052-X (print), 3-540-85053-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85053-3.

Omatu:2009:DCA

[ORB^{+}09] Sigeru Omatu, Miguel P. Rocha, José Bravo, Florentino Fernández, Emilio Corchado, Andrés Bustillo, and Juan M. Corchado, editors. Distributed Computing, Artificial Intelligence, Bioinformatics, Soft Computing, and Ambient Assisted Living: 10th International Work-Conference on Artificial Neural Networks, IWANN 2009 Workshops, Salamanca, Spain, June 10–12, 2009. Proceedings, Part II, volume 5518 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02480-7 (print), 3-642-02481-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02481-8.

Ozbudak:2012:AFF

[ÖRH12] Ferruh Özbudak and Francisco Rodríguez-Henríquez, editors. Arithmetic of Finite Fields: 4th International Workshop, WAIFI 2012, Bochum, Germany, July 16–19, 2012. Proceedings, volume 7369 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31661-1 (print), 3-642-31662-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31662-3.

OHare:2007:ESA

[OROD07] Gregory M. P. O’Hare, Alessandro Ricci, Michael J. O’Grady, and Oğuz Dikenelli, editors. Engineering Societies in the Agents World VII: 7th International Workshop, ESAW 2006 Dublin, Ireland, September 6– 8, 2006 Revised Selected and Invited Papers, volume 4457 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75522-5 (print), 3-540-75524-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75524-1.

Odijk:1989:PPAa

[ORS89a] Eddy Odijk, M. Rem, and J.-C. Syre, editors. PARLE’89: Parallel Architectures and Languages Europe: Eindhoven, the Netherlands, June 12–16, 1989: proceedings, volume 365 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51284-5 (USA: v. 1). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .P3184 1989 v. 1–2 (1989). US$36.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0365.htm;
 http://www.springerlink.
com/content/978-0-387-
51284-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=365.

Odijk:1989:PPAb

[ORS89b] Eddy Odijk, M. Rem, and J.-C. Syre, editors. PARLE’89: Parallel Architectures and Languages Europe: Eindhoven, the Netherlands, June 12–16, 1989: Proceedings. Part II. Parallel Languages, volume 366 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51285-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .P3184 1989 v. 1–2 (1989). US$36.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0366.

                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        51285-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=366.

Omondi:2003:ACS

[OS03a] Amos R. Omondi and Stanislav Sedukhin, editors. Advances in Computer Systems Architecture: 8th Asia-Pacific Conference, AC-SAC 2003, Aizu-Wakamatsu, Japan, September 23– 26, 2003: Proceedings, volume 2823 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20122-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A73 A28 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2823.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20122-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2823;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        13479.

OSullivan:2003:RAC

[O’S03b] B. O’Sullivan, editor. Recent advances in constraints: joint ERCIM/ CologNet International Workshop on Constraint Solving and Constraint Logic Programming, Cork, Ireland, June 19–21, 2002: Selected Papers, volume 2627 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00986-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .I63 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2627.htm;
 http://www.springerlink.
com/content/978-3-540-
00986-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2627.

Ohlbach:2004:PPS

[OS04] Hans Juergen Ohlbach and Sebastian Schaffert, editors. Principles and Practice of Semantic Web Reasoning: Second International Workshop, PPSWR 2004, St. Malo, France, September 6–10, 2004: Proceedings, volume 3208 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22961-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3208.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22961-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3208;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100228.

Okada:2007:ACS

[OS07] Mitsu Okada and Ichiro Satoh, editors. Advances in Computer Science — ASIAN 2006. Secure Software and Related Issues: 11th Asian Computing Science Conference, Tokyo, Japan, December 6–8, 2006, Revised Selected Papers, volume 4435 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77504-8 (print), 3-540-77505-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
77505-8.

Oliver:2009:IFO

[OS09] Miquel Oliver and Sebastià Sallent, editors. The Internet of the Future: 15th Open European Summer School and IFIP TC6.6 Workshop, EUNICE 2009, Barcelona, Spain, September 7–9, 2009. Proceedings, volume 5733 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03699-6 (print), 3-642-03700-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03700-9.

Onieva:2008:IST

[OSC^{+}08] Jose A. Onieva, Damien Sauveron, Serge Chaumette, Dieter Gollmann, and Konstantinos Markantonakis, editors. Information Security Theory and Practices. Smart Devices, Convergence and Next Generation Networks: Second IFIP WG 11.2 International Workshop, WISTP 2008, Seville, Spain, May 13– 16, 2008. Proceedings, volume 5019 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79965-6 (print), 3-540-79966-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79966-5.

ONeill:2002:AIC

[OSR^{+}02] Michael O’Neill, Richard F. E. Sutcliffe, Conor Ryan, Malachy Eaton, and Niall J. L. Griffith, editors. Artificial Intelligence and Cognitive Science: 13th Irish International Conference, AICS 2002, Limerick, Ireland, September 12–13, 2002. Proceedings, volume 2464 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44184-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I627 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2464.htm;
 http://www.springerlink.
com/content/978-3-540-
44184-7;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2464.

Overhage:2007:SAC

[OSRS07] Sven Overhage, Clemens A.

Szyperski, Ralf Reussner, and Judith A. Stafford, editors. Software Architectures, Components, and Applications: Third International Conference on Quality of Software Architectures, QoSA 2007, Medford, MA, USA, July 11–23, 2007, Revised Selected Papers, volume 4880 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77617-6 (print), 3-540-77619-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77619-2.

Ossowski:1999:COA

[Oss99] Sascha Ossowski. Coordination in artificial agent societies: social structures and its implications for autonomous problem-solving agents, volume 1535 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65495-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 221 pp. LCCN QA76.76.I58 O87 1999.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1535.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65495-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1535.

Owe:2011:FCT

[OST11] Olaf Owe, Martin Steffen, and Jan Arne Telle, editors. Fundamentals of Computation Theory: 18th International Symposium, FCT 2011, Oslo, Norway, August 22–25, 2011. Proceedings, volume 6914 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22952-2 (print), 3-642-22953-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22953-4.

Omicini:2011:DAL

[OSV11] Andrea Omicini, Sebastian Sardina, and Wamberto Vasconcelos, editors. Declarative Agent Languages and Technologies VIII: 8th International Workshop, DALT 2010, Toronto, Canada, May 10, 2010, Revised, Selected and Invited Papers, volume 6619 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20714-6 (print), 3-642-20715-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20715-0.

Orejas:2001:ALP

[OSvL01] F. (Fernando) Orejas, P. G. (Paul G.) Spirakis, and J. (Jan) van Leeuwen, editors. Automata, languages and programming: 28th international colloquium, ICALP 2001, Crete, Greece, July 8–12, 2001: Proceedings, volume 2076 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42287-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I551 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2076.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42287-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2076.

Orgun:2007:AAA

[OT07] Mehmet A. Orgun and John Thornton, editors. AI 2007: Advances in Artificial Intelligence: 20th Australian Joint Conference on Artificial Intelligence, Gold Coast, Australia, December 2–6, 2007. Proceedings, volume 4830 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76926-9 (print), 3-540-76928-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76928-6.

Ochmanski:2008:MFC

[OT08] Edward Ochmański and Jerzy Tyszkiewicz, editors. Mathematical Foundations of Computer Science 2008: 33rd International Symposium, MFCS 2008, Toruń, Poland, August 25–29, 2008. Proceedings, volume 5162 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85237-9 (print), 3-540-85238-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85238-4.

Ogihara:2011:TAM

[OT11] Mitsunori Ogihara and Jun Tarui, editors. Theory and Applications of Models of Computation: 8th Annual Conference, TAMC 2011, Tokyo, Japan, May 23–25, 2011. Proceedings, volume 6648 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20876-2 (print), 3-642-20877-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20877-5.

Ottmann:1987:ALP

[Ott87] Thomas Ottmann, editor. Automata, languages, and programming: 14th international colloquium, Karlsruhe, Federal Republic of Germany, July 13–17, 1987: proceedings, volume 267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18088-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.267. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0267.htm;
 http://www.springerlink.
com/content/978-0-387-
18088-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=267.

“ICALP 87”–Spine.

Omicini:2001:ESAa

[OTZ01] Andrea Omicini, Robert Tolksdorf, and Franco Zambonelli, editors. Engineering societies in the agents world: First International Workshop, ESAW 2000, Berlin, Germany, August 21, 2000: Revised Papers, volume 1972 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41477-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 E83 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1972.
htm;   http://www.springerlink.
com/content/978-3-540-
41477-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1972.

Ouaknine:2009:FMA

[OV09] Joël Ouaknine and Frits W.

Vaandrager, editors. Formal Modeling and Analysis of Timed Systems: 7th International Conference, FORMATS 2009, Budapest, Hungary, September 14– 16, 2009. Proceedings, volume 5813 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04367-4 (print), 3-642-04368-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04368-0.

Overmars:1983:DDD

[Ove83] Mark H. Overmars. The design of dynamic data structures, volume 156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-12330-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 181 pp. LCCN QA76.9.D35 O93 1983.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0156.
htm;   http://www.springerlink.
com/content/978-0-387-
12330-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=156.

ONeill:2008:GPE

[OVG^{+}08] Michael O’Neill, Leonardo Vanneschi, Steven Gustafson, Anna Isabel Esparcia Alcázar, Ivanoe De Falco, Antonio Della Cioppa, and Ernesto Tarantino, editors. Genetic Programming: 11th European Conference, EuroGP 2008, Naples, Italy, March 26–28, 2008. Proceedings, volume 4971 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78670-8 (print), 3-540-78671-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78671-9.

Okamoto:2007:PKC

[OW07] Tatsuaki Okamoto and Xiaoyun Wang, editors. Public Key Cryptography — PKC 2007: 10th International Conference on Practice and Theory in Public-Key Cryptography Beijing, China, April 16–20, 2007. Proceedings, volume 4450 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71676-9 (print), 3-540-71677-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71677-8.

Oliveira:2009:FMF

[OW09] Marcel Vinícius Medeiros Oliveira and Jim Woodcock, editors. Formal Methods: Foundations and Applications: 12th Brazilian Symposium on Formal Methods, SBMF 2009 Gramado, Brazil, August 19– 21, 2009 Revised Selected Papers, volume 5902 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10451-7 (print), 3-642-10452-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10452-7.

Oquendo:2004:SAF

[OWM04] Flavio Oquendo, Brian Warboys, and Ron Morrison, editors. Software Architecture: First European Workshop, EWSA 2004, St. Andrews, UK, May 21–22, 2004: Proceedings, volume 3047 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22000-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .E975 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3047.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22000-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3047;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97879.

Orlowska:2003:SOC

[OWPY03] Maria E. Orlowska, Sanjiva Weerawarana, Michael P. Papazoglou, and Jian Yang, editors. Service-Oriented Computing—ICSOC 2003: First International Conference, Trento, Italy, December 15–18, 2003: Proceedings, volume 2910 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20681-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I3224 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2910.
htm;   http://www.springerlink.
com/content/978-3-540-
20681-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2910;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94513.

Orchard:2004:IAA

[OYA04] Bob Orchard, Chunsheng Yang, and Moonis Ali, editors. Innovations in Applied Artificial Intelligence: 17th International Conference on Industrial and Engineering Applications of Artificial Intelligence and Expert Systems, IEA/AIE 2004, Ottawa, Canada, May 17–20, 2004: Proceedings, volume 3029 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22007-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .L43x v.3029. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3029.
htm;   http://www.springerlink.
com/content/978-3-540-
22007-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3029;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97304.

Okadome:2007:PCQ

[OYM07] Takeshi Okadome, Tatsuya Yamazaki, and Mounir Makhtari, editors. Pervasive Computing for Quality of Life Enhancement: 5th International Conference On Smart Homes and Health Telematics, ICOST 2007, Nara, Japan, June 21–23, 2007. Proceedings, volume 4541 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73034-6 (print), 3-540-73035-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73035-4.

Olive:2003:ACM

[OYY03] Antoni Olive, Masatoshi Yoshikawa, and Eric S. K. Yu, editors. Advanced Conceptual Modeling Techniques: ER 2002 workshops-ECDM, MobiMod, IWCMQ, and eCOMO, Tampere, Finland, October 7–11, 2002: Proceedings, volume 2784 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20255-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26 A38 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2784.
htm;   http://www.springerlink.
com/content/978-3-540-
20255-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2784;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b12013.

Osterby:1983:DMS

[OZ83] O. Osterby and Zahari Zlatev. Direct methods for sparse matrices, volume 157 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-12676-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 127 pp. LCCN QA188 .O87 1983. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0157.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        12676-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=157.

Oliveira:2001:FFM

[OZ01] Jose N. Oliveira and Pamela Zave, editors. FME 2001: formal methods for increasing software productivity: [10th] International Symposium of Formal Methods Europe, Berlin, Germany, March 12–16, 2001: Proceedings, volume 2021 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41791-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I593 2001; QA267.A1 L43 no.2021. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2021.htm;
 http://www.springerlink.
com/content/978-3-540-
41791-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2021.

Ozok:2009:OCS

[OZ09] A. Ant Ozok and Panayiotis Zaphiris, editors. Online Communities and Social Computing: Third International Conference, OCSC 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009. Proceedings, volume 5621 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02773-3 (print), 3-642-02774-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02774-1.

Ozok:2011:OCS

[OZ11] A. Ant Ozok and Panayiotis Zaphiris, editors. Online Communities and Social Computing: 4th International Conference, OCSC 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011. Proceedings, volume 6778 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21795-8 (print), 3-642-21796-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21796-8.

Pande:2001:COS

[PA01] Santosh Pande and D. P. Agrawal, editors. Compiler Optimizations for Scalable Parallel Systems: Languages, Compilation Techniques, and Run Time Systems, volume 1808 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41945-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xxvii + 778 pp. LCCN QA76.76.C65 C665 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1808.htm;
 http://www.springerlink.
com/content/978-3-540-
41945-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1808.

Pena:2003:IFL

[PA03a] Ricardo Pena and Thomas Arts, editors. Implementation of functional languages: 14th International Workshop, IFL 2002, Madrid, Spain, September 16–18, 2002: Revised Papers, volume 2670 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40190-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .I58 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2670.htm;
 http://www.springerlink.
com/content/978-3-540-
40190-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2670.

Pires:2003:PAI

[PA03b] Fernando Moura Pires and Salvador Abreu, editors. Progress in Artificial Intelligence: 11th Portuguese Title: Conference on Artificial Intelligence, EPIA 2003, Beja, Portugal, December 4–7, 2003: Proceedings, volume 2902 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20589-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .P67 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2902.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20589-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2902;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b94425.

Parashar:2009:DCI

[PA09] Manish Parashar and Sanjeev K. Aggarwal, editors. Distributed Computing and Internet Technology: 5th International Conference, ICD-CIT 2008 New Delhi, India, December 10– 12, 2008. Proceedings, volume 5375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-89736-4 (print), 3-540-89737-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89737-8.

Pach:2005:GDI

[Pac05] János Pach, editor.

Graph drawing: 12th International Symposium, GD 2004, New York, NY, USA, September 29– October 2, 2004, Revised selected papers, volume 3383 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24528-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 2004.

URL http://www.springerlink.
com/content/978-3-540-
24528-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3383;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105810.

Pasko:2008:HOM

[PAC08] Alexander Pasko, Valery Adzhiev, and Peter Comninos, editors. Heterogeneous Objects Modelling and Applications: Collection of Papers on Foundations and Practice, volume 4889 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68441-7 (print), 3-540-68443-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68443-5.

Padget:1999:CBH

[Pad99] Julian A. Padget. Collaboration between human and artificial societies: coordination and agentbased distributed computing, volume 1624 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66930-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 300 pp. LCCN QA76.9.D5 C615 1999.

                                        URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t1624.
htm;   http://www.springerlink.
com/content/978-3-540-
66930-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1624.

Pan:2006:TLD

[PAD^{+}06] Zhigeng Pan, Ruth Aylett, Holger Diener, Xiaogang Jin, Stefan Göbel, and Li Li, editors. Technologies for E-Learning and Digital Entertainment: First International Conference, Edutainment 2006, Hangzhou, China, April 16–19, 2006. Proceedings, volume 3942 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33423-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33423-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3942.

Palma:2008:HPC

[PAD^{+}08a] José M. Laginha M.

Palma, Patrick R. Amestoy, Michel Daydé, Marta Mattoso, and João Correia Lopes, editors. High Performance Computing for Computational Science — VECPAR 2008: 8th International Conference, Toulouse, France, June 24–27, 2008. Revised Selected Papers, volume 5336 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92858-8 (print), 3-540-92859-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92859-1.

Pavlou:2008:MCM

[PAD08b] George Pavlou, Toufik Ahmed, and Tasos Dagiuklas, editors. Management of Converged Multimedia Networks and Services: 11th IFIP/IEEE International Conference on Management of Multimedia and Mobile Networks and Services, MMNS 2008, Samos Island, Greece, September 22–26, 2008. Proceedings, volume 5274 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87358-9 (print), 3-540-87359-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87359-4.

Page:1998:ILP

[Pag98] David Page, editor. Inductive logic programming: 8th International Conference, ILP-98, Madison, Wisconsin, USA, July 22–24, 1998: Proceedings, volume 1446 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64738-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I52 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1446.htm;
http://www.springerlink.
com/content/978-3-540-
64738-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1446.

Paiva:2000:AIT

[Pai00]  Ana Paiva.  Affec-

tive interactions: towards a new generation of computer interfaces, volume 1814 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41520-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 234 pp. LCCN TK7887.5 .A38 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1814.
htm;   http://www.springerlink.
com/content/978-3-540-
41520-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1814.

Paige:2009:TPM

[Pai09] Richard F. Paige, editor. Theory and Practice of Model Transformations: Second International Conference, ICMT 2009, Zurich, Switzerland, June 29–30, 2009. Proceedings, volume 5563 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02407-6 (print), 3-642-02408-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02408-5.

Palamidessi:2000:CCT

[Pal00a] Catuscia Palamidessi, editor. CONCUR 2000— Concurrency Theory: 11th International Conference, University Park, PA, USA, August 22– 25, 2000. Proceedings, volume 1877 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67897-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1877.htm;
http://www.springerlink.
com/content/978-3-540-
67897-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1877.

Palsberg:2000:SAI

[Pal00b] Jens Palsberg, editor.

Static analysis: 7th International Symposium, SAS 2000, Santa Barbara, CA, USA, June 29–July 1, 2000: Proceedings, volume 1824 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67668-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1824. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1824.htm;
 http://www.springerlink.
com/content/978-3-540-
67668-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1824.

Palazzo:2001:ETI

[Pal01] Sergio Palazzo, editor. Evolutionary trends of the Internet: 2001 Thyrrhenian International Workshop on Digital Communications, IWDC 2001, Taormina, Italy, September 17– 20, 2001: Proceedings, volume 2170 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42592-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2170; TK5103.7 .T48 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2170.htm;
 http://www.springerlink.
com/content/978-3-540-
42592-2;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2170.

Also available via the World Wide Web.

Palamidessi:2003:LPI

[Pal03] Catuscia Palamidessi, editor. Logic Programming: 19th International Conference, ICLP 2003, Mumbai, India, December 9–13, 2003: Proceedings, volume 2916 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20642-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2916.htm;
http://www.springerlink.
com/content/978-3-540-
20642-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2916;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94619.

Palsberg:2009:SAS

[Pal09] Jens Palsberg, editor. Semantics and Algebraic Specification: Essays Dedicated to Peter D. Mosses on the Occasion of His 60th Birthday, volume 5700 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04163-9 (print), 3-642-04164-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04164-8.

Pan:1984:HMM

[Pan84] Victor Pan. How to multiply matrices faster, volume 179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CODEN LNCSD9. ISBN 0-387-13866-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 212 pp. LCCN QA188 .P36 1984. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0179.htm;
 http://www.springerlink.
com/content/978-0-387-
13866-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=179.

Papazoglou:1995:OOO

[Pap95] M. Papazoglou, editor. OOER’95: objectoriented and entityrelationship modelling: 14th International Conference, Gold Coast, Australia, December 13– 15, 1995: proceedings, volume 1021 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60672-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .O54 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1021.htm;
 http://www.springerlink.
com/content/978-3-540-
60672-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1021.

Proceedings of the fourteenth International Conference on Object-Oriented and Entity-Relationship Modelling.

Paredaens:1984:ALP

[Par84] Jan Paredaens, editor.

Automata, languages, and programming: 11th Colloquium, Antwerp, Belgium, July 16–20, 1984, volume 172 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-13345-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA267.A1 L43 no.172. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0172.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        13345-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=172.

Proceedings of the 11th Colloquium on Automata, Languages, and Programming.

Parikh:1985:LPB

[Par85] Rohit Parikh, editor. Logics of programs: Brooklyn, June 17–19, 1985: proceedings, volume 193 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15648-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L58551 1985. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0193.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        15648-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=193.

Papers presented at a conference held at Brooklyn College June 17–19,

1985.

ParisiPresice:1998:RTA

[Par98] Francesco Parisi Presice, editor. Recent trends in algebraic development techniques: 12th International Workshop, WADT ’97, Tarquinia, Italy, June 3–7, 1997: Selected Papers, volume 1376 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64299-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47W33 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1376.
htm;   http://www.springerlink.
com/content/978-3-540-
64299-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1376.

Parashar:2002:GCG

[Par02] Manish Parashar, editor. Grid computing

— GRID 2002: Third International Workshop, Baltimore, MD, USA, November 18, 2002: Proceedings, volume 2536 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00133-6 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 G74 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2536.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00133-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2536.

Also available via the

World Wide Web.

Parker:2009:CCI

[Par09] Matthew G. Parker, editor. Cryptography and Coding: 12th IMA International Conference, Cryptography and Coding 2009, Cirencester, UK, December 15–17, 2009. Proceedings, volume 5921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10867-9 (print), 3-642-10868-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10868-6.

Pasetti:2002:SFE

[Pas02] Alessandro Pasetti. Software Frameworks and Embedded Control Systems, volume 2231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43189-6. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 293 pp. LCCN QA76.9.A73 P375 2002; TK7895.E42 P37 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2231.htm;
 http://www.springerlink.
com/content/978-3-540-
43189-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2231.

Thesis (doctoral) — Universität, Konstanz, 2001.

Pasareanu:2009:MCS

[Păs09] Corina S. Păsăreanu, editor. Model Checking Software: 16th International SPIN Workshop, Grenoble, France, June 26–28, 2009. Proceedings, volume 5578 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02651-6 (print), 3-642-02652-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-

                                        02652-2.

Paterson:1990:ALP

[Pat90] Michael S. Paterson, editor. Automata, languages, and programming: 17th international colloquium, Warwick University, England, July 16–20, 1990: proceedings, volume 443 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52826-1 (Berlin), 0-387-52826-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.443. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0443.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        52826-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=443.

ICALP (International Colloquium on Automata, Languages, and Programming) was sponsored by the European Association for Theoretical Computer Science.

Paterson:2000:AEA

[Pat00] Michael S. Paterson, editor. Algorithms— ESA 2000: 8th Annual European Symposium, Saarbrücken, Germany, September 5–8, 2000. Proceedings, volume 1879 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41004-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1879.htm;
 http://www.springerlink.
com/content/978-3-540-
41004-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1879.

Paterno:2002:HCI

[Pat02] Fabio Paternò, editor.

Human computer interaction with mobile devices: 4th International Symposium, Mobile HCI 2002, Pisa, Italy, September 18– 20, 2002: Proceedings, volume 2411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44189-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.59 .M67 2002. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2411.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44189-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2411.

Also available via the

World Wide Web.

Paterson:2003:CCI

[Pat03] Kenneth G. Paterson, editor. Cryptography and Coding: 9th IMA International Conference, Cirencester, UK, December 16–18, 2003: Proceedings, volume 2898 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20663-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .C76 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2898.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20663-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2898;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b93924.

Paterson:2011:ACE

[Pat11] Kenneth G. Paterson, editor. Advances in Cryptology — EUROCRYPT 2011: 30th Annual International Conference on the Theory and Applications of Cryptographic Techniques, Tallinn, Estonia, May 15–19, 2011. Proceedings, volume 6632 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. ISBN 3-642-20464-3 (soft cover), 3-642-20465-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20464-7;   http://www.
springerlink.com/content/
978-3-642-20465-4.

Paulisch:1993:DES

[Pau93] Frances Newbery Paulisch.

The design of an extendible [sic] graph editor, volume 704 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57090-X (Berlin), 0-387-57090-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 184 pp. LCCN T385 .P376 1993. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0704.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        57090-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=704.

Revision of the author’s thesis (doctoral — University of Karlsruhe, 1991).

Paulson:1994:IGT

[Pau94] Lawrence C. Paulson. Isabelle: a generic theorem prover, volume 828 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58244-4, 0-387-58244-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xvii + 321 pp. LCCN QA76.9.A96 P38 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0828.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58244-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=828.

Pauli:2001:LBR

[Pau01] Josef Pauli. Learning-Based Robot Vision: Principles and Applications, volume 2048 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42108-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 288 pp. LCCN TJ211.3 .P38 2001; QA267.A1 L43 no.2048. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2048.htm;
 http://www.springerlink.
com/content/978-3-540-
42108-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2048.

Paun:2003:MCI

[Pau03] Gheorghe Paun, editor. Membrane computing: International Workshop, WMC-CdeA 2002: Curtea-de-Arges, Romania, August 19– 23, 2002: Revised Papers, volume 2597 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00611-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.887 .W53 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2597.htm;
 http://www.springerlink.
com/content/978-3-540-
00611-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2597.

Padget:2010:COI

[PAV^{+}10] Julian Padget, Alexander Artikis, Wamberto Vasconcelos, Kostas Stathis, Viviane Torres da Silva, Eric Matson, and Axel Polleres, editors. Coordination, Organizations, Institutions and Norms in Agent Systems V: COIN 2009 International Workshops. COIN@AAMAS 2009, Budapest, Hungary, May 2009, COIN@IJCAI 2009, Pasadena, USA, July 2009, COIN@MALLOW 2009, Turin, Italy, September 2009. Revised Selected Papers, volume 6069 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14961-8 (print), 3-642-14962-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14962-7.

Pazienza:1997:IEM

[Paz97] Maria Teresa Pazienza, editor. Information extraction: a multidisciplinary approach to an emerging information technology: international summer school, SCIE-97, Frascati, Italy, July 14–18-97, volume 1299 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63438-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.T48S38 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1299.htm;
 http://www.springerlink.
com/content/978-3-540-
63438-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1299.

Pazienza:1999:IET

[Paz99] Maria Teresa Pazienza, editor. Information Extraction: Towards Scalable, Adaptable Systems, volume 1714 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66625-7. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 164 pp. LCCN QA76.9.T48 I64 1999. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t1714.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66625-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1714.

Pazienza:2003:IEW

[Paz03] Maria Teresa Pazienza, editor. Information Extraction in the Web Era: Natural Language Communication for Knowledge Acquisition and Intelligent Information Agents, volume 2700 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40579-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.I58 I55 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2700.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40579-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2700;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/

b11781. Proceedings of SCIE 2002: Summer Convention on Information Extraction.

Pustejovsky:1992:LSK

[PB92] J. Pustejovsky and S. Bergler, editors. Lexical semantics and knowledge representation: first SIGLEX Workshop, Berkeley, CA, USA, June 13, 1991: proceedings, volume 627 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 0-387-55801-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN P325 .S53 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0627.htm;
 http://www.springerlink.
com/content/978-0-387-
55801-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=627.

Plaza:1997:KAM

[PB97] Enric Plaza and Richard Benjamins, editors. Knowledge acquisition, modelling and management: 10th European Workshop, EKAW ’97, Sant Feliu de Guixols, Catalonia, Spain, October 15–18, 1997: proceedings, volume 1319 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63592-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E92 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1319.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63592-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1319.

Paschke:2007:ARI

[PB07] Adrian Paschke and Yevgen Biletskiy, editors. Advances in Rule Interchange and Applications: International Symposium, RuleML 2007, Orlando, Florida, October 25–26, 2007. Proceedings, volume 4824 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75974-3 (print), 3-540-75975-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75975-1.

Peter:2008:AEH

[PB08a] Christian Peter and Russell Beale, editors. Affect and Emotion in Human-Computer Interaction: From Theory to Applications, volume 4868 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85098-8 (print), 3-540-85099-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85099-1.

Pirhonen:2008:HAI

[PB08b] Antti Pirhonen and Stephen Brewster, editors. Haptic and Audio Interaction Design: Third International Workshop, HAID 2008 Jyväskylä, Finland, September 15–16, 2008 Proceedings, volume 5270 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87882-3 (print), 3-540-87883-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87883-4.

Pfeifer:2009:WWM

[PB09] Tom Pfeifer and Paolo Bellavista, editors. Wired-Wireless Multimedia Networks and Services Management: 12th IFIP/ IEEE International Conference on Management of Multimedia and Mobile Networks and Services, MMNS 2009, Venice, Italy, October 26–27, 2009. Proceedings, volume 5842 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04993-1 (print), 3-642-04994-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04994-1.

Piloty:1983:CR

[PBB^{+}83] Robert Piloty, M. Barbacci, D. Borrione, D. Dietmeyer, E. Hill, and R. Skelly. CONLAN report, volume 151 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 3-540-12275-3, 0-387-12275-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 174 pp. LCCN QA76.73.C28 C65 1983. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t0151.htm;
http://www.springerlink.
com/content/978-3-540-
12275-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=151.

Pal:2005:PRM

[PBB05] Sankar K. Pal, Sanghamitra Bandyopadhyay, and Sambhunath Biswas, editors. Pattern recognition and machine intelligence: first international conference, PReMI 2005, Kolkata, India, December 20–22, 2005: proceedings, volume 3776 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30506-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q327 .I58 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30506-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3776.

Polani:2004:RRS

[PBBY04] Daniel Polani, Brett Browning, Andrea Bonarini, and Kazuo Yoshida, editors. RoboCup 2003: Robot Soccer World Cup VII, volume 3020 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22443-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3020.htm;
 http://www.springerlink.
com/content/978-3-540-
22443-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3020;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98623.

Pezzulo:2008:CAU

[PBCF08] Giovanni Pezzulo, Martin V. Butz, Cristiano Castelfranchi, and Rino Falcone, editors. The Challenge of Anticipation: A Unifying Framework for the Analysis and Design of Artificial Cognitive Systems, volume 5225 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87701-0 (print), 3-540-87702-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
87702-8.

Priami:2010:TCS

[PBG^{+}10] Corrado Priami, Rainer Breitling, David Gilbert, Monika Heiner, and Adelinde M. Uhrmacher, editors. Transactions on Computational Systems Biology XII: Special Issue on Modeling Methodologies, volume 5945 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11711-2 (print), 3-642-11712-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .T73 2010. URL http:

//www.springerlink.
com/content/978-3-642-
11712-1.

Peters:2002:ECL

[PBGK02] C. Peters, M. Braschler, J. Gonzalo, and M. Kluck, editors. Evaluation of cross-language information retrieval systems: Second Workshop of the Cross-Language Evaluation Forum, CLEF 2001, Darmstadt, Germany, September 3–4, 2001: Revised Papers, volume 2406 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44042-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z667.5.C75 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2406.
htm;   http://www.springerlink.
com/content/978-3-540-
44042-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2406.

Peters:2003:ACL

[PBGK03] Carol Peters, Martin Braschler, Julio Gonzalo, and Michael Kluck, editors. Advances in Cross-Language Information Retrieval: Third Workshop of the Cross-Language Evaluation Forum, CLEF 2002: Revised Papers, volume 2785 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40830-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z667.5 .C75 2003. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2785.
htm;   http://www.springerlink.
com/content/978-3-540-
40830-7;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2785;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12018.

Perner:2003:MDA

[PBH03] Petra Perner, Rüdiger Brause, and Hermann-Georg Holzhütter, editors. Medical Data Analysis: 4th International Symposium, IS-MDA 2003, Berlin, Germany, October 9–10, 2002: Proceedings, volume 2868 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20282-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R858.A2. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2868.htm;
http://www.springerlink.
com/content/978-3-540-
20282-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2868;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b14289.

Pierre:2003:AHM

[PBK03] Samuel Pierre, Michel Barbeau, and Evangelos Kranakis, editors. Ad-Hoc, Mobile, and Wireless Networks: Second International Conference, ADHOC-NOW 2003, Montréal, Canada, October 8–10, 2003: Proceedings, volume 2865 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20260-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK6570.M6 A34 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2865.
htm;   http://www.springerlink.
com/content/978-3-540-
20260-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2865;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b14230.

Pardo:2010:CPP

[PBK^{+}10] Thiago Alexandre Salgueiro Pardo, António Branco, Aldebaro Klautau, Renata Vieira, and Vera Lúcia Strube de Lima, editors. Computational Processing of the Portuguese Language: 9th International Conference, PROPOR 2010, Porto Alegre, RS, Brazil, April 27–30, 2010. Proceedings, volume 6001 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12319-8 (print), 3-642-12320-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12320-7.

Priami:2009:TCS

[PBP09] Corrado Priami, Ralph-Johan Back, and Ion Petre, editors. Transactions on Computational Systems Biology XI, volume 5750 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04185-X (print), 3-642-04186-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04186-0.

Priami:2011:TCS

[PBPdV11] Corrado Priami, Ralph-Johan Back, Ion Petre, and Erik de Vink, editors. Transactions on Computational Systems Biology XIII, volume 6575 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19747-7 (print), 3-642-19748-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19748-2.

Popescu-Belis:2008:MLMa

[PBRB08] Andrei Popescu-Belis, Steve Renals, and Hervé Bourlard, editors. Machine Learning for Multimodal Interaction: 4th International Workshop, MLMI 2007, Brno, Czech Republic, June 28–30, 2007, Revised Selected Papers, volume 4892 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78154-4 (print), 3-540-78155-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78155-4.

Popescu-Belis:2008:MLMb

[PBS08] Andrei Popescu-Belis and Rainer Stiefelhagen, editors. Machine Learning for Multimodal Interaction: 5th International Workshop, MLMI 2008, Utrecht, The Netherlands, September 8– 10, 2008. Proceedings, volume 5237 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85852-0 (print), 3-540-85853-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85853-9.

Pezzulo:2009:ABA

[PBSB09] Giovanni Pezzulo, Martin V. Butz, Olivier Sigaud, and Gianluca Baldassarre, editors. Anticipatory Behavior in Adaptive Learning Systems: From Psychological Theories to Artificial Cognitive Systems, volume 5499 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02564-1 (print), 3-642-02565-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        02565-5.

Proll:2005:CWT

[PBW05] Birgit Pröll, Kurt Bauknecht, and Hannes Werthner, editors. E-commerce and Web Technologies: 6th International Conference, EC-Web 2005, Copenhagen, Denmark, August 23–26, 2005: Proceedings, volume 3590 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28467-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28467-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3590.

Paterson:1999:CPM

[PC99] Michael S. Paterson and Maxime Crochemore, editors. Combinatorial pattern matching: 10th annual symposium, CPM 99, Warwick University, UK, July 22–24, 1999: Proceedings, volume 1645 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66278-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1645.
htm;   http://www.springerlink.
com/content/978-3-540-
66278-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1645.

Pontelli:2000:PAD

[PC00] Enrico Pontelli and Vitor Santos Costa, editors. Practical Aspects of Declarative Languages: Second International Workshop, PADL 2000, Boston, MA, USA, January 17–18, 2000. Proceedings, volume 1753 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-66992-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1753.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1753.
htm;   http://www.springerlink.
com/content/978-3-540-
66992-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1753.

Park:2005:ISC

[PC05] Choonsik Park and Seongtaek Chee, editors. Information security and cryptology: ICISC 2004: 7th international conference, Seoul, Korea, December 2–3, 2004. Revised selected papers, volume 3506 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26226-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26226-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3506;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b137120.

Priss:2002:CSI

[PCA02] Uta Priss, Dan Corbett, and Galia Angelova, editors. Conceptual structures: integration and interfaces: 10th International Conference on Conceptual Structures, ICCS 2002, Borovets, Bulgaria, July 15–19, 2002: Proceedings, volume 2393 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43901-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2 .I58 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2393.htm;
 http://www.springerlink.
com/content/978-3-540-
43901-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2393.

Park:2009:AIS

[PCA^{+}09] Jong Hyuk Park, Hsiao-Hwa Chen, Mohammed Atiquzzaman, Changhoon Lee, Tai hoon Kim, and Sang-Soo Yeo, editors. Advances in Information Security and Assurance: Third International Conference and Workshops, ISA 2009, Seoul, Korea, June 25–27, 2009. Proceedings, volume 5576 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02616-8 (print), 3-642-02617-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02617-1.

Priami:2006:TCSa

[PCE06] Corrado Priami, Luca Cardelli, and Stephen Emmott, editors. Transactions on Computational Systems Biology IV, volume 3939 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33245-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .T735 2006b. URL http:

                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        book%26isbn=978-3-540-
                                        33245-9;   http://www.
                                        springerlink.com/content/
                                        978-3-540-33245-9;
                                        http://www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3939.

Pan:2005:PDP

[PCG^{+}05a] Yi Pan, Daoxu Chen, Minyi Guo, Jiannong Cao, and Jack J. Dongarra, editors. Parallel and distributed processing and applications: third international symposium, ISPA 2005, Nanjing, China, November 2–5, 2005: proceedings, volume 3758 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29769-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I88 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29769-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3758.

Peters:2005:MIA

[PCG^{+}05b] Carol Peters, Paul Clough, Julio Gonzalo, Gareth J. F. Jones, Michael Kluck, and Bernardo Magnini, editors. Multilingual Information Access for Text, Speech and Images: 5th Workshop of the Cross-Language Evaluation Forum, CLEF 2004, Bath, UK, September 15–17, 2004, Revised Selected Papers, volume 3491 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27420-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z667.5 .C75 2004 852. URL

http://www.springerlink.
com/content/978-3-540-
27420-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3491.

Peters:2007:EMM

[PCG^{+}07] Carol Peters, Paul Clough, Fredric C. Gey, Jussi Karlgren, Bernardo Magnini, Douglas W. Oard, Maarten de Rijke, and Maximilian Stempfhuber, editors. Evaluation of Multilingual and Multi-modal Information Retrieval: 7th Workshop of the Cross-Language Evaluation Forum, CLEF 2006, Alicante, Spain, September 20–22, 2006, Revised Selected Papers, volume 4730 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74998-5 (print), 3-540-74999-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74999-8.

Peters:2010:MIAb

[PCG^{+}10] Carol Peters, Barbara Caputo, Julio Gonzalo, Gareth J. F. Jones, Jayashree Kalpathy-Cramer, Henning Müller, and Theodora Tsikrika, editors. Multilingual Information Access Evaluation II. Multimedia Experiments: 10th Workshop of the Cross-Language Evaluation Forum, CLEF 2009, Corfu, Greece, September 30 — October 2, 2009, Revised Selected Papers, volume 6242 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15750-5 (print), 3-642-15751-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15751-6.

Pan:2006:AAR

[PCH^{+}06] Zhigeng Pan, Adrian Cheok, Michael Haller, Rynson W. H. Lau, Hideo Saito, and Ronghua Liang, editors. Advances in Artificial Reality and Tele-Existence: 16th International Conference on Artificial Reality and Telexistence, ICAT 2006, Hangzhou, China, November 29 — December 1, 2006. Proceedings, volume 4282 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49776-5 (print), 3-540-49779-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49779-0.

Pan:2012:SWJ

[PCK^{+}12] Jeff Z. Pan, Huajun Chen, Hong-Gee Kim, Juanzi Li, Zhe Wu, Ian Horrocks, Riichiro Mizoguchi, and Zhaohui Wu, editors. The Semantic Web: Joint International Semantic Technology Conference, JIST 2011, Hangzhou, China, December 4–7, 2011. Proceedings, volume 7185 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29922-9 (print), 3-642-29923-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29923-0.

Pan:2010:TEI

[PCM^{+}10] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, Xiaopeng Zhang, and Kevin Wong, editors. Transactions on Edutainment IV, volume 6250 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14483-7 (print), 3-642-14484-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14484-4.

Pan:2011:TEVb

[PCM11] Zhigeng Pan, Adrian David Cheok, and Wolfgang Müller, editors. Transactions on Edutainment VI, volume 6758 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22638-8 (print), 3-642-22639-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22639-7.

Pan:2012:TEVa

[PCM^{+}12a] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, Maiga Chang, and Mingmin Zhang, editors. Transactions on Edutainment VII, volume 7145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29049-3 (print), 3-642-29050-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29050-3.

Pan:2012:TEVb

[PCM^{+}12b] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, Maiga Chang, and Mingmin Zhang, editors. Transactions on Edutainment VIII, volume 7220 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31438-4 (print), 3-642-31439-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31439-1.

Pan:2009:TEIb

[PCMC09] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, and Maiga Chang, editors. Transactions on Edutainment III, volume 5940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-11244-7 (print), 3-642-11245-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-642-
11245-4.

Pan:2008:TE

[PCME08] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, and Abdennour El Rhalibi, editors. Transactions on Edutainment I, volume 5080 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69737-3 (print), 3-540-69744-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69744-2.

Pan:2009:TEIa

[PCME09] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, and Abdennour El Rhalibi, editors. Transactions on Edutainment II, volume 5660 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03269-9 (print), 3-642-03270-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03270-7.

Pan:2011:TEVa

[PCMY11] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, and Xubo Yang, editors. Transactions on Edutainment V, volume 6530 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18451-0 (print), 3-642-18452-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18452-9.

Pan:2010:CCIa

[PCN10a] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Computational Collective Intelligence. Technologies and Applications: Second International Conference, ICCCI 2010, Kaohsiung, Taiwan, November 10– 12, 2010. Proceedings, Part I, volume 6421 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16692-X (print), 3-642-16693-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-642-
16693-8.

Pan:2010:CCIb

[PCN10b] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Computational Collective Intelligence. Technologies and Applications: Second International Conference, ICCCI 2010, Kaohsiung, Taiwan, November 10– 12, 2010, Proceedings, Part II, volume 6422 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16731-4 (print), 3-642-16732-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16732-4.

Pan:2010:CCIc

[PCN10c] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Computational Collective Intelligence. Technologies and Applications: Second International Conference, ICCCI 2010, Kaohsiung, Taiwan, November 10– 12, 2010. Proceedings, Part III, volume 6423 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16695-4 (print), 3-642-16696-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16696-9.

Pan:2012:IIDa

[PCN12a] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Intelligent Information and Database Systems: 4th Asian Conference, ACI-IDS 2012, Kaohsiung, Taiwan, March 19–21, 2012, Proceedings, Part I, volume 7196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28486-8 (print), 3-642-28487-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28487-8.

Pan:2012:IIDb

[PCN12b] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Intelligent Information and Database Systems: 4th Asian Conference, ACI-IDS 2012, Kaohsiung, Taiwan, March 19–21, 2012, Proceedings, Part II, volume 7197 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28489-2 (print), 3-642-28490-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28490-8.

Pan:2012:IIDc

[PCN12c] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Intelligent Information and Database Systems: 4th Asian Conference, ACI-IDS 2012, Kaohsiung, Taiwan, March 19–21, 2012, Proceedings, Part III, volume 7198 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28492-2 (print), 3-642-28493-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28493-9.

Perales:2003:PRI

[PCPS03] Francisco José Perales, Aurélio J. C. Campilho, Nicolás Pérez, and Alberto Sanfeliu, editors. Pattern Recognition and Image Analysis: First Iberian Conference, IbPRIA 2003, Puerto de Andratx, Mallorca, Spain, June 4–6, 2003: Proceedings, volume 2652 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40217-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1650 .I27 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2652.htm;
 http://www.springerlink.
com/content/978-3-540-
40217-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2652.

Perrott:2007:HPC

[PCS^{+}07] Ronald Perrott, Barbara M. Chapman, Jaspal Subhlok, Rodrigo Fernandes de Mello, and Laurence T. Yang, editors. High Performance Computing and Communications: Third International Conference, HPCC 2007, Houston, USA, September 26– 28, 2007. Proceedings, volume 4782 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75443-1 (print), 3-540-75444-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75444-2.

Perrin:1996:DPP

[PD96] Guy-Rene Perrin and Alain Darte. The data parallel programming model: foundations, HPF realization, and scientific applications, volume 1132 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61736-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 284 pp. LCCN QA76.642 .D38 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1132.htm;
http://www.springerlink.
com/content/978-3-540-
61736-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1132.

Perales:2004:AMD

[PD04] Francisco J. Perales and Bruce A. Draper, editors. Articulated Motion and Deformable Objects: Third International Workshop, AMDO 2004, Palma de Mallorca, Spain, September 22–24, 2004, Proceedings, volume 3179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22958-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3179.htm;
 http://www.springerlink.
com/content/978-3-540-
22958-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3179;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100217.

Polleres:2011:RWS

[PdA^{+}11] Axel Polleres, Claudia d’Amato, Marcelo Arenas, Siegfried Handschuh, Paula Kroner, Sascha Ossowski, and Peter Patel-Schneider, editors. Reasoning Web. Semantic Technologies for the Web of Data: 7th International Summer School 2011, Galway, Ireland, August 23– 27, 2011, Tutorial Lectures, volume 6848 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23031-8 (print), 3-642-23032-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23032-5.

Priami:2008:TCSb

[PDAN08] Corrado Priami, Falko Dressler, Ozgur B. Akan, and Alioune Ngom, editors. Transactions on Computational Systems Biology X, volume 5410 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92272-5 (print), 3-540-92273-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92273-5.

Peters:2009:ESM

[PDF^{+}09] Carol Peters, Thomas Deselaers, Nicola Ferro, Julio Gonzalo, Gareth J. F. Jones, Mikko Kurimo, Thomas Mandl, Anselmo Peñas, and Vivien Petras, editors. Evaluating Systems for Multilingual and Multimodal Information Access: 9th Workshop of the Cross-Language Evaluation Forum, CLEF 2008, Aarhus, Denmark, September 17–19, 2008, Revised Selected Papers, volume 5706 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04446-8 (print), 3-642-04447-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04447-2.

Pirotte:1992:ADT

[PDG92] A. Pirotte, C. Delobel, and G. Gottlob, editors. Advances in database technology — EDBT ’92: 3rd International Conference on Extending Database Technology, Vienna, Austria, March 23–27, 1992: proceedings, volume 580 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55270-7 (Berlin), 0-387-55270-7 (New York).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55844 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0580.htm;
 http://www.springerlink.
com/content/978-3-540-
55270-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=580.

Spine title: EDBT ’92.

Palma:1999:VPP

[PDH99] José M. L. M. Palma, J. J. Dongarra, and Vicente Hernández, editors. Vector and parallel processing — VECPAR ’98: Third International Conference, Porto, Portugal, June 21–23, 1998: selected papers and invited talks, volume 1573 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66228-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1573. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1573.htm;
 http://www.springerlink.
com/content/978-3-540-
66228-0;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1573.

Palma:2001:VPP

[PDH01] José M. L. M. Palma, Jack J. Dongarra, and Vicente Hernández, editors. Vector and parallel processing — VECPAR 2000: 4th International Conference, Porto, Portugal, June 21–23, 2000: selected papers and invited talks, volume 1981 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41999-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I552 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1981.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41999-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1981.

Palma:2003:HPC

[PDHS03] José M. L. M. Palma, Jack Dongarra, Vicente Hernández, and A. Augusto Sousa, editors. High performance computing for computational science, VECPAR 2002: 5th International Conference, Porto, Portugal, June 26–28, 2002: selected papers and invited talks, volume 2565 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00852-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I552 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2565.htm;
 http://www.springerlink.
com/content/978-3-540-
00852-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2565.

Peters:2010:MIAa

[PDK^{+}10] Carol Peters, Giorgio Maria Di Nunzio, Mikko Kurimo, Thomas Mandl, Djamel Mostefa, Anselmo Peñas, and Giovanna Roda, editors. Multilingual Information Access Evaluation I. Text Retrieval Experiments: 10th Workshop of the Cross-Language Evaluation Forum, CLEF 2009, Corfu, Greece, September 30 — October 2, 2009, Revised Selected Papers, volume 6241 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15753-X (print), 3-642-15754-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15754-7.

Pavon:2012:AAI

[PDMFF12] Juan Pavón, Néstor D. Duque-Méndez, and Rubén Fuentes-Fernández, editors. Advances in Artificial Intelligence — IB-ERAMIA 2012: 13th Ibero–American Conference on AI, Cartagena de Indias, Colombia, November 13–16, 2012. Proceedings, volume 7637 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34653-7 (print), 3-642-34654-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34654-5.

Palma:2011:HPC

[PDML11] José M. Laginha M. Palma, Michel Daydé, Osni Marques, and João Correia Lopes, editors. High Performance Computing for Computational Science — VEC-PAR 2010: 9th International conference, Berkeley, CA, USA, June 22–25, 2010, Revised Selected Papers, volume 6449 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19327-7 (print), 3-642-19328-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19328-6.

Paterno:2012:AIT

[PdRM^{+}12] Fabio Paternò, Boris de Ruyter, Panos Markopoulos, Carmen Santoro, Evert van Loenen, and Kris Luyten, editors. Ambient Intelligence: Third International Joint Conference, AmI 2012, Pisa, Italy, November 13– 15, 2012. Proceedings, volume 7683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34897-1 (print), 3-642-34898-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        34898-3.

Panigrahi:2010:SEM

[PDSD10] Bijaya Ketan Panigrahi, Swagatam Das, Ponnuthurai Nagaratnam Suganthan, and Subhransu Sekhar Dash, editors. Swarm, Evolutionary, and Memetic Computing: First International Conference on Swarm, Evolutionary, and Memetic Computing, SEMCCO 2010, Chennai, India, December 16–18, 2010. Proceedings, volume 6466 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17562-7 (print), 3-642-17563-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17563-3.

Panigrahi:2012:SEM

[PDSN12] Bijaya Ketan Panigrahi, Swagatam Das, Ponnuthurai Nagaratnam Suganthan, and Pradipta Kumar Nanda, editors. Swarm, Evolutionary, and Memetic Computing: Third International Conference, SEMCCO 2012, Bhubaneswar, India, December 20–22, 2012. Proceedings, volume 7677 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35379-7 (print), 3-642-35380-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35380-2.

Pelc:2007:DCI

[Pel07] Andrzej Pelc, editor.

Distributed Computing: 21st International Symposium, DISC 2007, Lemesos, Cyprus, September 24–26, 2007. Proceedings, volume 4731 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75141-6 (print), 3-540-75142-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75142-7.

Peleg:2011:DCI

[Pel11] David Peleg, editor.

Distributed Computing: 25th International Symposium, DISC 2011, Rome, Italy, September 20–22, 2011. Proceedings, volume 6950 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24099-2 (print), 3-642-24100-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24100-0.

Perner:2001:MLD

[Per01] Petra Perner, editor. Machine learning and data mining in pattern recognition: Second International Workshop, MLDM 2001, Leipzig, Germany, July 25–27, 2001: Proceedings, volume 2123 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42359-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2123; Q327 .M56 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2123.htm;
 http://www.springerlink.
com/content/978-3-540-
42359-1;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2123.

Perner:2002:ADM

[Per02a] Petra Perner, editor. Advances in Data Mining: Applications in E-Commerce, Medicine, and Knowledge Management, volume 2394 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44116-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 A334 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2394.htm;
http://www.springerlink.
com/content/978-3-540-
44116-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2394.

All papers were presented at the second Industrial Conference on Data Mining ICDM 2002 in Leipzig.

Perner:2002:DMM

[Per02b]  Petra Perner.  Data

Mining on Multimedia Data, volume 2558 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00317-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 131 pp. LCCN QA76.575 .P45 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2558.
htm;   http://www.springerlink.
com/content/978-3-540-
00317-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2558.

Also available via the World Wide Web.

Perner:2005:ADM

[Per05] Petra Perner, editor. Advances in data mining: applications in image mining, medicine and biotechnology, management and environmental control, and telecommunications; 4th industrial conference on data mining, ICDM 2004, Leipzig, Germany, July 4–7, 2004, revised selected papers, volume 3275 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24054-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 I53 2004. URL http:

//www.springerlink.

com/content/978-3-540-
24054-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3275;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104334.

Perner:2006:ADM

[Per06] Petra Perner, editor. Advances in data mining: applications in medicine, Web mining, marketing, image and signal mining: 6th Industrial Conference on Data Mining, ICDM 2006, Leipzig, Germany, July 14–15, 2006: proceedings, volume 4065 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-36036-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343. URL

http://www.springerlink.
com/content/978-3-540-
36036-0.

Perner:2007:ADM

[Per07a] Petra Perner, editor. Advances in Data Mining. Theoretical Aspects and Applications: 7th Industrial Conference, ICDM 2007, Leipzig, Germany, July 14–18, 2007. Proceedings, volume 4597 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73434-1 (print), 3-540-73435-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73435-2.

Perner:2007:MLD

[Per07b] Petra Perner, editor. Machine Learning and Data Mining in Pattern Recognition: 5th International Conference, MLDM 2007, Leipzig, Germany, July 18–20, 2007. Proceedings, volume 4571 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73498-8 (print), 3-540-73499-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73499-4.

Perner:2008:ADM

[Per08] Petra Perner, editor. Advances in Data Mining. Medical Applications, E-Commerce, Marketing, and Theoretical Aspects: 8th Industrial Conference, ICDM 2008 Leipzig, Germany, July 16–18, 2008 Proceedings, volume 5077 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70717-4 (print), 3-540-70720-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70720-2.

Perner:2009:ADM

[Per09a] Petra Perner, editor. Advances in Data Mining. Applications and Theoretical Aspects: 9th Industrial Conference, ICDM 2009, Leipzig, Germany, July 20 — 22, 2009. Proceedings, volume 5633 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03066-1 (print), 3-642-03067-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03067-3.

Perner:2009:MLD

[Per09b] Petra Perner, editor.

Machine Learning and Data Mining in Pattern Recognition: 6th International Conference, MLDM 2009, Leipzig, Germany, July 23–25, 2009. Proceedings, volume 5632 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03069-6 (print), 3-642-03070-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03070-3.

Perner:2010:ADM

[Per10a] Petra Perner, editor. Advances in Data Mining. Applications and Theoretical Aspects: 10th Industrial Conference, ICDM 2010, Berlin, Germany, July 12–14, 2010. Proceedings, volume 6171 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14399-7 (print), 3-642-14400-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14400-4.

Pernici:2010:AIS

[Per10b] Barbara Pernici, editor.

Advanced Information Systems Engineering: 22nd International Conference, CAiSE 2010, Hammamet, Tunisia, June 7–9, 2010. Proceedings, volume 6051 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13093-3 (print), 3-642-13094-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13094-6.

Perner:2011:ADM

[Per11a] Petra Perner, editor. Advances in Data Mining. Applications and Theoretical Aspects: 11th Industrial Conference, ICDM 2011, New York, NY, USA, August 30 — September 3, 2011. Proceedings, volume 6870 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23183-7 (print), 3-642-23184-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23184-1.

Perner:2011:MLD

[Per11b] Petra Perner, editor. Machine Learning and Data Mining in Pattern Recognition: 7th International Conference, MLDM 2011, New York, NY, USA, August 30 — September 3, 2011. Proceedings, volume 6871 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23198-5 (print), 3-642-23199-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23199-5.

Persiano:2011:AGT

[Per11c] Giuseppe Persiano, editor. Algorithmic Game Theory: 4th International Symposium, SAGT 2011, Amalfi, Italy, October 17–19, 2011. Proceedings, volume 6982 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24828-4 (print), 3-642-24829-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24829-0.

Perner:2012:ADM

[Per12a] Petra Perner, editor. Advances in Data Mining. Applications and Theoretical Aspects: 12th Industrial Conference, ICDM 2012, Berlin, Germany, July 13–20, 2012. Proceedings, volume 7377 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31487-2 (print), 3-642-31488-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31488-9.

Perner:2012:MLD

[Per12b] Petra Perner, editor.

Machine Learning and Data Mining in Pattern Recognition: 8th International Conference, MLDM 2012, Berlin, Germany, July 13–20, 2012. Proceedings, volume 7376 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31536-4 (print), 3-642-31537-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31537-4.

Peterson:1980:CPS

[Pet80] James Lyle Peterson. Computer programs for spelling correction: an experiment in program design, volume 96 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10259-0. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 213 pp. LCCN QA76.6 .P474. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0096.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        10259-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=96.

Pettorossi:1992:MPL

[Pet92] A. Pettorossi, editor. Meta-programmming in logic: Third International Workshop, META-92, Uppsala, Sweden, June 10–12, 1992: proceedings, volume 649 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-56282-6 (Berlin), 0-387-56282-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73 .I567 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0649.htm;
 http://www.springerlink.
com/content/978-3-540-
56282-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=649.

Pettersson:1999:CNS

[Pet99] Mikael Pettersson. Compiling natural semantics, volume 1549 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65968-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 240 pp. LCCN QA76.7 .P493 1999. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1549.
htm;   http://www.springerlink.
com/content/978-3-540-
65968-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1549.

Peters:2001:CLI

[Pet01] C. Peters, editor. Crosslanguage information retrieval and evaluation: workshop of the Cross-Language Evaluation Forum, CLEF 2000, Lisbon, Portugal, September 21–22, 2000: Revised Papers, volume 2069 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42446-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z667.5 .C75 2000; QA267.A1 L43 no.2069. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2069.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42446-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2069.

Pettorossi:2002:LBP

[Pet02] A. Pettorossi, editor. Logic Based Program Synthesis and Transformation: 11th International Workshop, LOPSTR 2001, Paphos, Cyprus, November 28– 30, 2001. Selected Papers, volume 2372 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43915-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I57 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2372.htm;
 http://www.springerlink.
com/content/978-3-540-
43915-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2372.

Petitcolas:2003:IHI

[Pet03] Fabien A. P. Petitcolas, editor. Information hiding: 5th International Workshop, IH 2002, Noordwijkerhout, the Netherlands, October 7–9, 2002: Revised Papers, volume 2578 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00421-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I49 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2578.htm;
 http://www.springerlink.
com/content/978-3-540-
00421-9;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2578.

Also available via the

World Wide Web.

Pezze:2003:FAS

[Pez03] Mauro Pezzè, editor. Fundamental approaches to software engineering: 6th International Conference, FASE 2003, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2003, Warsaw, Poland, April 7–11, 2003: Proceedings, volume 2621 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00899-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.75 .F44 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2621.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00899-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2621.

Also available via the

World Wide Web.

Pastor:2005:AIS

[PF05] Oscar Pastor and João Falcão e Cunha, editors. Advanced information systems engineering: 17th international conference, CAISE 2005, Porto, Portugal, June 13–17, 2005, proceedings, volume 3520 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26095-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .C76 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26095-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3520;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136788.

Perales:2006:AMD

[PF06] Francisco J. Perales and Robert B. Fisher, editors. Articulated Motion and Deformable Objects: 4th International Conference, AMDO 2006, Port d’Andratx, Mallorca, Spain, July 11– 14, 2006. Proceedings, volume 4069 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36031-X (print), 3-540-36032-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        36032-2.

Preparata:2007:FAF

[PF07a] Franco P. Preparata and Qizhi Fang, editors. Frontiers in Algorithmics: First Annual International Workshop, FAW 2007, Lanzhou, China, August 1–3, 2007. Proceedings, volume 4613 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73813-4 (print), 3-540-73814-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73814-5.

Prodan:2007:GCE

[PF07b] Radu Prodan and Thomas Fahringer, editors. Grid Computing: Experiment Management, Tool Integration, and Scientific Workflows, volume 4340 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CO-DEN LNCSD9. ISBN 3-540-69261-4 (print), 3-540-69262-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 295 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69262-1.

Perales:2008:AMD

[PF08] Francisco J. Perales and Robert B. Fisher, editors. Articulated Motion and Deformable Objects: 5th International Conference, AMDO 2008, Port d’Andratx, Mallorca, Spain, July 9– 11, 2008. Proceedings, volume 5098 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70516-3 (print), 3-540-70517-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70517-8.

Perales:2010:AMD

[PF10] Francisco J. Perales and Robert B. Fisher, editors. Articulated Motion and Deformable Objects: 6th International Conference, AMDO 2010, Port d’Andratx, Mallorca, Spain, July 7– 9, 2010. Proceedings, volume 6169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14060-2 (print), 3-642-14061-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14061-7.

Paragios:2005:VGL

[PFCS05] Nikos Paragios, Olivier Faugeras, Tony Chan, and Christoph Schnoerr, editors. Variational, geometric, and level set methods in computer vision: third international workshop, VLSM 2005, Beijing, China, October 16, 2005: proceedings, volume 3752 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29348-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .I329 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29348-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3752.

Patt:2010:HPE

[PFD^{+}10] Yale N. Patt, Pierfrancesco Foglia, Evelyn Duesterwald, Paolo Faraboschi, and Xavier Martorell, editors. High Performance Embedded Architectures and Compilers: 5th International Conference, HiPEAC 2010, Pisa, Italy, January 25–27, 2010. Proceedings, volume 5952 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11514-4 (print), 3-642-11515-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11515-8.

Pfenning:1994:LPA

[Pfe94] Frank Pfenning, editor. Logic programming and automated reasoning: 5th International Conference, LPAR ’94, Kiev, Ukraine, July 16–22, 1994: proceedings, volume 822 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58216-9 (Berlin), 0-387-58216-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 I565 1994. DM72.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0822.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        58216-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=822.

Pfenning:2006:TRA

[Pfe06] Frank Pfenning, editor. Term Rewriting and Applications: 17th International Conference, RTA 2006 Seattle, WA, USA, August 12–14, 2006 Proceedings, volume 4098 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36834-5 (print), 3-540-36835-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        36835-9.

Pfenning:2007:ADC

[Pfe07] Frank Pfenning, editor. Automated Deduction — CADE-21: 21st International Conference on Automated Deduction Bremen, Germany, July 17–20, 2007 Proceedings, volume 4603 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73594-1 (print), 3-540-73595-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73595-3.

Pfitzmann:1996:DSS

[Pfi96] Birgit Pfitzmann. Digital signature schemes: general framework and fail-stop signatures, volume 1100 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61517-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 396 pp. LCCN QA76.9.A25 P444 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1100.
htm;   http://www.springerlink.
com/content/978-3-540-
61517-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1100.

Revision of thesis (Ph.

D.)–University of Hildes-

heim, 1993.

Pfitzmann:2000:IHT

[Pfi00] Andreas Pfitzmann, editor. Information hiding: Third International Workshop, IH’99, Dresden, Germany, September 29–October 1, 1999: Proceedings, volume 1768 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67182-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1768.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1768.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67182-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1768.

Pfitzmann:2001:ACE

[Pfi01] Birgit Pfitzmann, editor. Advances in cryptology: EUROCRYPT 2001: International Conference on the Theory and Application of Cryptographic Techniques, Innsbruck, Austria, May 6–10, 2001: Proceedings, volume 2045 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, April 25, 2001. CODEN LNCSD9. ISBN 3-540-42070-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 2001; QA76.9.A25 E96 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2045.
htm;   http://link.
springer.de/link/service/
series/0558/tocs/t2045.
htm;   http://www.springerlink.
com/content/978-3-540-
42070-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2045.

Pflanz:2002:LED

[Pfl02] Matthias Pflanz. Online Error Detection and Fast Recover Techniques for Dependable Embedded Processors, volume 2270 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43318-X. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 126 pp. LCCN TK7895.E42 P455 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2270.htm;
 http://www.springerlink.
com/content/978-3-540-
43318-7;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2270.

Perales:2012:AMD

[PFM12] Francisco J. Perales, Robert B. Fisher, and Thomas B. Moeslund, editors. Articulated Motion and Deformable Objects: 7th International Conference, AMDO 2012, Port d’Andratx, Mallorca, Spain, July 11– 13, 2012. Proceedings, volume 7378 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31566-6 (print), 3-642-31567-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31567-1.

Pierre:2001:MAT

[PG01] Samuel Pierre and Roch Glitho, editors. Mobile agents for telecommunication applications: Third International Workshop, MATA 2001, Montréal, Canada, August 14– 16, 2001: Proceedings, volume 2164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42460-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 I585 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2164.htm;
http://www.springerlink.
com/content/978-3-540-
42460-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2164.

Prakash:2009:ISS

[PG09] Atul Prakash and Indranil Sen Gupta, editors. Information Systems Security: 5th International Conference, ICISS 2009 Kolkata, India, December 14– 18, 2009 Proceedings, volume 5905 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10771-0 (print), 3-642-10772-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10772-6.

Panayiotopoulos:2005:IVA

[PGA^{+}05] Themis Panayiotopoulos, Jonathan Gratch, Ruth Aylett, Daniel Ballin, Patrick Olivier, and Thomas Rist, editors. Intelligent virtual agents: 5th international working conference, IVA 2005, Kos, Greece, September 12–14, 2005: proceedings, volume 3661 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28738-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65 I93 2005.

URL http://www.springerlink.
com/content/978-3-540-
28738-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3661.

Peters:2004:CEM

[PGBK04] Carol Peters, Julio Gonzalo, Martin Braschler, and Michael Kluck, editors. Comparative evaluation of multilingual information access systems: 4th workshop of the cross-language evaluation forum, CLEF 2003, Trondheim, Norway, August 21–22, 2003, revised papers, volume 3237 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-24017-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
24017-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3237.

Pieprzyk:2007:ISP

[PGD07] Josef Pieprzyk, Hossein Ghodosi, and Ed Dawson, editors. Information Security and Privacy: 12th Australasian Conference, ACISP 2007, Townsville, Australia, July 2–4, 2007. Proceedings, volume 4586 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73457-0 (print), 3-540-73458-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73458-1.

Peters:2006:AMI

[PGG^{+}06] Carol Peters, Fredric C.

Gey, Julio Gonzalo, Henning Müller, Gareth J. F. Jones, Michael Kluck, Bernardo Magnini, and Maarten de Rijke, editors. Accessing Multilingual Information Repositories: 6th Workshop of the Cross-Language Evaluation Forum, CLEF 2005, Vienna, Austria, 21–23 September, 2005, Revised Selected Papers, volume 4022 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45697-X (print), 3-540-45700-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45700-8.

Palamidessi:1998:PDP

[PGM98] Catuscia Palamidessi, Hugh Glaser, and K. Meinke, editors. Principles of declarative programming: 10th International Symposium, PLILP’98: held jointly with the 6th International Conference, ALP’98, Pisa, Italy, September 16–18, 1998: Proceedings, volume 1490 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65012-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1490. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1490.htm;

http://www.springerlink.
com/content/978-3-540-
65012-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1490.

Preneel:1993:CSI

[PGV93] Bart Preneel, Rene Govaerts, and J. Vandewalle, editors. Computer security and industrial cryptography: state of the art and evolution: ESAT course, Leuven, Belgium, May 21– 23, 1991, volume 741 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57341-0 (U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C6373 1993. DM58.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0741.htm;
http://www.springerlink.
com/content/978-0-387-
57341-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=741.

Pelillo:1997:EMM

[PH97] Marcello Pelillo and Edwin R. Hancock, editors. Energy minimization methods in computer vision and pattern recognition: International Workshop EMM-CVPR’97, Venice, Italy, May 21–23, 1997: proceedings, volume 1223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62909-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E54 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1223.htm;
 http://www.springerlink.
com/content/978-3-540-
62909-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1223.

Papers presented at the International Workshop on Energy Minimization Methods in Computer Vision and Pattern Recognition, held at the University of Venice, Italy.

Perales:2002:AMD

[PH02] Francisco Jose Perales and Edwin R. Hancock, editors. Articulated Motion and Deformable Objects: Second International Workshop, AMDO 2002, Palma de Mallorca, Spain, November 21–23, 2002. Proceedings, volume 2492 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00149-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .A486 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2492.
htm;   http://www.springerlink.
com/content/978-3-540-
00149-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2492.

Papatriantafilou:2004:PDS

[PH04] Marina Papatriantafilou and Philippe Hunel, editors. Principles of Distributed Systems: 7th International Conference, OPODIS 2003, La Martinique, French West Indies, December 10–13, 2003, Revised Selected Papers, volume 3144 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22667-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3144.htm;
 http://www.springerlink.
com/content/978-3-540-
22667-3;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3144;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99477.

Pinho:2006:RST

[PH06] Luís Miguel Pinho and Michael González Harbour, editors. Reliable Software Technologies — Ada-Europe 2006: 11th Ada-Europe International Conference on Reliable Software Technologies, Porto, Portugal, June 5–9, 2006. Proceedings, volume 4006 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34663-5 (print), 3-540-34664-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34664-7.

Paul:2010:GTC

[PH10] Christophe Paul and Michel Habib, editors. Graph-Theoretic Concepts in Computer Science: 35th International Workshop, WG 2009, Montpellier, France, June 24–26, 2009. Revised Papers, volume 5911 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11408-3 (print), 3-642-11409-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11409-0.

Parampalli:2011:ISP

[PH11a] Udaya Parampalli and Philip Hawkes, editors. Information Security and Privacy: 16th Australasian Conference, ACISP 2011, Melbourne, Australia, July 11–13, 2011. Proceedings, volume 6812 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22496-2 (print), 3-642-22497-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22497-3.

Pelillo:2011:SBP

[PH11b] Marcello Pelillo and Edwin R. Hancock, editors. Similarity-Based Pattern Recognition: First International Workshop, SIMBAD 2011, Venice, Italy, September 28– 30, 2011. Proceedings, volume 7005 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24470-X (print), 3-642-24471-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24471-1.

Picco:2012:WSN

[PH12] Gian Pietro Picco and Wendi Heinzelman, editors. Wireless Sensor Networks: 9th European Conference, EWSN 2012, Trento, Italy, February 15–17, 2012. Proceedings, volume 7158 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28168-0 (print), 3-642-28169-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28169-3.

Pfahringer:2010:DSI

[PHH10] Bernhard Pfahringer, Geoff Holmes, and Achim Hoffmann, editors. Discovery Science: 13th International Conference, DS 2010, Canberra, Australia, October 6–8, 2010. Proceedings, volume 6332 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16183-9 (print), 3-642-16184-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16184-1.

Palade:2003:KBIb

[PHJ03a] Vasile Palade, Bob Howlett, and Lakhmi Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 7th International Conference, KES 2003, Oxford, UK, September 3–5, 2003: Proceedings, volume 2774 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40804-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2774.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40804-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2774;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12003.

Palade:2003:KBIa

[PHJ03b] Vasile Palade, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based Intelligent Information and Engineering Systems: 7th International Conference, KES 2003, Oxford, UK, September 3–5, 2003, volume 2773 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40803-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7801 .I5155 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2773.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40803-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2773;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12002.

Ploug:2010:PTI

[PHOK10] Thomas Ploug, Per Hasle, and Harri Oinas-Kukkonen, editors. Persuasive Technology: 5th International Conference, PERSUASIVE 2010, Copenhagen, Denmark, June 7–10, 2010. Proceedings, volume 6137 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13225-1 (print), 3-642-13226-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13226-1.

Priami:2006:TCSb

[PHPL06] Corrado Priami, Xiaohua Hu, Yi Pan, and Tsau Young Lin, editors. Transactions on Computational Systems Biology V, volume 4070 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36048-4 (print), 3-540-36049-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
36049-0.

Paige:2009:MDA

[PHR09] Richard F. Paige, Alan Hartman, and Arend Rensink, editors. Model Driven Architecture — Foundations and Applications: 5th European Conference, ECMDA-FA 2009, Enschede, The Netherlands, June 23– 26, 2009. Proceedings, volume 5562 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02673-7 (print), 3-642-02674-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02674-4.

Ponce:2006:TCL

[PHSZ06] Jean Ponce, Martial Hebert, Cordelia Schmid, and Andrew Zisserman, editors. Toward Category-Level Object Recognition, volume 4170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68794-7 (print), 3-540-68795-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68795-5.

Page:2011:TFP

[PHZ11] Rex Page, Zoltán Horváth, and Viktória Zsók, editors. Trends in Functional Programming: 11th International Symposium, TFP 2010, Norman, OK, USA, May 17–19, 2010. Revised Selected Papers, volume 6546 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22940-9 (print), 3-642-22941-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22941-1.

Perner:2005:MLD

[PI05] Petra Perner and Atsushi Imiya, editors. Machine Learning and Data Mining in Pattern Recognition: 4th International Conference, MLDM 2005, Leipzig, Germany, July 9–11, 2005: Proceedings, volume 3587 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26923-1, 3-540-26923-

1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q327 .M56 1999. URL http:

//www.springerlink.
com/content/978-3-540-
26923-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3587.

Pichler:1986:ACE

[Pic86] Franz Pichler, editor. Advances in cryptology: Eurocrypt 85: proceedings of a workshop on the theory and application of cryptographic techniques, Linz, Austria, April, 1985, volume 219 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16468-5 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E961 1985. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0219.htm;
 http://www.springerlink.
com/content/978-0-387-
16468-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

0302-9743&volume=219.

“The workshop was sponsored by International Association for Cryptologic Research … [et al.]”–T.p. verso.

Picco:2001:MAI

[Pic01] Gian Pietro Picco, editor. Mobile Agents: 5th International Conference, MA 2001 Atlanta, GA, USA, December 2–4, 2001. Proceedings, volume 2240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42952-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M3 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2240.htm;
http://www.springerlink.
com/content/978-3-540-
42952-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2240.

Pierson:2005:DMG

[Pie05] Jean-Marc Pierson, editor. Data management in grids: first VLDB Workshop, DMG 2005, Trondheim, Norway, September 2–3, 2005: revised selected papers, volume 3836 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-31212-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 D38 2005.

URL http://www.springerlink.
com/content/978-3-540-
31212-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3836.

Pieprzyk:2008:ACA

[Pie08] Josef Pieprzyk, editor. Advances in Cryptology

— ASIACRYPT 2008: 14th International Conference on the Theory and Application of Cryptology and Information Security, Melbourne, Australia, December 7–11, 2008. Proceedings, volume 5350 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89254-0 (print), 3-540-89255-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89255-7.

Pieprzyk:2010:TCC

[Pie10] Josef Pieprzyk, editor.

Topics in Cryptology — CT-RSA 2010: The Cryptographers’ Track at the RSA Conference 2010, San Francisco, CA, USA, March 1–5, 2010. Proceedings, volume 5985 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11924-7 (print), 3-642-11925-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11925-5.

Priami:2006:TCSd

[PIMN06] Corrado Priami, Anna Ingólfsdóttir, Bud Mishra, and Hanne Riis Nielson, editors. Transactions on Computational Systems Biology VII, volume 4230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48837-5 (print), 3-540-48839-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48839-2.

Pin:1989:FPF

[Pin89] Jean Eric Pin, editor. Formal properties of finite automata and applications: LITP Spring School on Theoretical Computer Science, Ramatuelle, France, May 23–27, 1988: proceedings, volume 386 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51631-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.S4 E26 1988. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0386.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51631-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=386.

Pingali:1995:LCP

[Pin95] Keshav Pingali, editor. Languages and compilers for parallel computing: 7th International Workshop, Ithaca, NY, USA, August 8–10, 1994: proceedings, volume 892 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-58868-X (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0892.htm;
 http://www.springerlink.
com/content/978-3-540-
58868-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=892.

“Revised version of papers presented at Seventh Annual Workshop on Languages and Compilers for Parallel Computers.”.

Prasanna:2005:DCS

[PISW05] Viktor K. Prasanna, Sitharama Iyengar, Paul G. Spirakis, and Matt Welsh, editors. Distributed Computing in Sensor Systems: First IEEE International Conference, DCOSS 2005, Marina del Rey, CA, USA, June 30 — July 1, 2005. Proceedings, volume 3560 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26422-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7872.D48 D36 2005.

URL http://www.springerlink.

                                        com/content/978-3-540-
                                        26422-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3560.

Pitt:1986:CTC

[Pit86] David Pitt, editor. Category theory and computer programming: tutorial and workshop, Guildford, UK, September 16–20, 1985: proceedings, volume 240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-17162-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .C3871 1986. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0240.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        17162-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=240.

Pitt:1989:CTC

[Pit89] David Pitt, editor. Category theory and computer science: Manchester, UK, September 5–8, 1989, proceedings, volume 389 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51662-X (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .C38821 1989. DM52.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0389.htm;
 http://www.springerlink.
com/content/978-0-387-
51662-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=389.

Papers presented at the 3rd in a series of biennial conferences.

Pitt:1991:CTC

[Pit91] David Pitt, editor. Category theory and computer science: Paris, France, September 3–6, 1991: proceedings, volume 530 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54495-X (Berlin), 0-387-54495-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S86 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0530.htm;
 http://www.springerlink.
com/content/978-3-540-
54495-1;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=530.

“Papers … presented at the fourth biennial Summer Conference on Category Theory and Computer Science, held in Paris, September 3–6, 1991”–Pref.

Plasil:1997:STP

[PJ97] Frantisek Plasil and Keith (Keith G.) Jeffery, editors. SOFSEM’97: theory and practice of informatics: 24th Seminar on Current Trends in Theory and Practice of Informatics, Milovy, Czech Republic, November 22–29, 1997:proceedings, volume 1338 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63774-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751.S62 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1338.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63774-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1338.

Pointcheval:2012:ACE

[PJ12] David Pointcheval and Thomas Johansson, editors. Advances in cryptology — EUROCRYPT 2012: 31st Annual International Conference on the Theory and Applications of Cryptographic Techniques, Cambridge, UK, April 15–19, 2012. Proceedings, volume 7237 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-29010-8, 3-642-29011-6 (e-book). LCCN QA76.9.A25 E97 2012.

URL http://www.springerlink.
com/content/978-3-642-
29010-7.

Pallis:2012:SOC

[PJC^{+}12] George Pallis, Mohamed Jmaiel, Anis Charfi, Sven Graupner, Yücel Karabulut, Sam Guinea, Florian Rosenberg, Quan Z. Sheng, Cesare Pautasso, and Sonia Ben Mokhtar, editors. Service-Oriented Computing — ICSOC 2011 Workshops: IC-SOC 2011, International Workshops WE-SOA, NFPSLAM-SOC, and Satellite Events, Paphos, Cyprus, December 5–8, 2011. Revised Selected Papers, volume 7221 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31874-6 (print), 3-642-31875-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31875-7.

Peters:2008:AMM

[PJM^{+}08] Carol Peters, Valentin Jijkoun, Thomas Mandl, Henning Müller, Douglas W. Oard, Anselmo Peñas, Vivien Petras, and Diana Santos, editors. Advances in Multilingual and Multimodal Information Retrieval: 8th Workshop of the Cross-Language Evaluation Forum, CLEF 2007, Budapest, Hungary, September 19–21, 2007, Revised Selected Papers, volume 5152 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85759-1 (print), 3-540-85760-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85760-0.

Plodereder:2000:RST

[PK00] Erhard Plodereder and H. B. Keller, editors. Reliable software technologies, Ada-Europe 2000: 5th Ada-Europe International Conference, Potsdam, Germany, June 26–30, 2000: Proceedings, volume 1845 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67669-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1845. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1845.htm;
 http://www.springerlink.
com/content/978-3-540-
67669-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1845.

Petitcolas:2003:DWF

[PK03] Fabien A. P. Petitcolas and Hyoung Joong Kim, editors. Digital watermarking: First International Workshop, IWDW 2002, Seoul, Korea, November 21–22, 2002: Revised Papers, volume 2613. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-01217-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I939 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2613.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        01217-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2613.

Also available via the

World Wide Web.

Pietka:2012:ITB

[PK12] Ewa Pietka and Jacek Kawa, editors. Information Technologies in Biomedicine: Third International Conference, ITIB 2012, Gliwice, Poland, June 11– 13, 2012. Proceedings, volume 7339 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31195-4 (print), 3-642-31196-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31196-3.

Pal:2006:DCI

[PKKG06] Ajit Pal, Ajay D. Kshemkalyani, Rajeev Kumar, and Arobinda Gupta, editors. Distributed Computing — IWDC 2005: 7th International Workshop, Kharagpur, India, December 27–30, 2005. Proceedings, volume 3741 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30959-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30959-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3741.

Pal:2004:NIP

[PKM04] Nikhil R. Pal, Nikola Kasabov, and Rajani K. Mudi, editors. Neural Information Processing: 11th International Conference, ICONIP 2004, Calcutta, India, November 22–25, 2004: Proceedings, volume 3316 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23931-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t3316.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23931-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3316;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b103766.

Pandurangan:2010:WAS

[PKM^{+}10] Gopal Pandurangan, V. S. Anil Kumar, Gu Ming, Yunhao Liu, and Yingshu Li, editors. Wireless Algorithms, Systems, and Applications: 5th International Conference, WASA 2010, Beijing, China, August 15–17, 2010. Proceedings, volume 6221 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14653-8 (print), 3-642-14654-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14654-1.

Paliouras:2001:MLA

[PKS01] Georgios Paliouras, Vangelis Karkaletsis, and Constantine D. Spyropoulos, editors. Machine Learning and Its Applications: Advanced Lectures, volume 2049 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42490-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 324 pp. LCCN QA325.5 .M33 2001; QA267.A1 L43 no.2049. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2049.htm;
 http://www.springerlink.
com/content/978-3-540-
42490-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2049.

Plaice:2002:DCW

[PKSS02] John Plaice, Peter G.

Kropf, Peter Schulthess, and Jacob Slonim, editors. Distributed Communities on the Web: 4th International Workshop, DCW 2002, Sydney, Australia, April 3– 5, 2002. Revised Papers, volume 2468 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00301-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D375 2002. URL http:

                                        //link.springer.de/
                                        link/service/series/
                                        0558/tocs/t2468.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00301-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2468.

Also available via the

World Wide Web.

Plattner:1996:BCN

[Pla96] B. (Bernhard) Plattner, editor. Broadband communications: networks, services, applications, future directions: 1996 International Zürich Seminar on Digital Communications, IZS ’96, Zürich, Switzerland, February 21–23, 1996: proceedings, volume 1044 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60895-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.4 .I59 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1044.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60895-0;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1044.

Peleg:2011:AIM

[PLC11] Mor Peleg, Nada Lavrač, and Carlo Combi, editors. Artificial Intelligence in Medicine: 13th Conference on Artificial Intelligence in Medicine, AIME 2011, Bled, Slovenia, July 2–6, 2011. Proceedings, volume 6747 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22217-X (print), 3-642-22218-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22218-4.

Pluim:2006:BIR

[PLG06] Josien P. W. Pluim, Boštjan Likar, and Frans A. Gerritsen, editors. Biomedical Image Registration: Third International Workshop, WBIR 2006, Utrecht, The Netherlands, July 9–11, 2006. Proceedings, volume 4057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35648-7 (print), 3-540-35649-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35649-3.

Prendinger:2008:IVA

[PLI08] Helmut Prendinger, James Lester, and Mitsuru Ishizuka, editors. Intelligent Virtual Agents: 8th International Conference, IVA 2008, Tokyo, Japan, September 1– 3, 2008. Proceedings, volume 5208 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85482-7 (print), 3-540-85483-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85483-8.

Popescu:2012:AWB

[PLK^{+}12] Elvira Popescu, Qing Li, Ralf Klamma, Howard Leung, and Marcus Specht, editors. Advances in Web-Based Learning — ICWL 2012: 11th International Conference, Sinaia, Romania, September 2– 4, 2012. Proceedings, volume 7558 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33641-8 (print), 3-642-33642-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33642-3.

Prescott:2012:BBS

[PLMV12] Tony J. Prescott, Nathan F.

Lepora, Anna Mura, and Paul F. M. J. Verschure, editors. Biomimetic and Biohybrid Systems: First International Conference, Living Machines 2012, Barcelona, Spain, July 9–12, 2012. Proceedings, volume 7375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31524-0 (print), 3-642-31525-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31525-1.

Plumer:1990:TPL

[Plu90] Lutz Plumer. Termination proofs for logic programs, volume 446 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52837-7 (Berlin), 0-387-52837-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 142 pp. LCCN QA76.63 .P58 1990. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0446.
htm;   http://www.springerlink.
com/content/978-3-540-
52837-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=446.

Rev. version of author’s thesis.

Perram:1996:DSA

[PM96] John W. Perram and Jean-Pierre Muller, editors. Distributed software agents and applications: 6th European Workshop on Modelling Autonomous Agents in a Multi-Agent World, MAAMAW’94, Odense, Denmark, August 3– 5, 1994: proceedings, volume 1069 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61157-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q337 .E87 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1069.htm;
 http://www.springerlink.
com/content/978-3-540-
61157-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1069.

Pajdla:2004:CVEa

[PM04a] Tomás Pajdla and Jirà Matas, editors. Computer Vision — ECCV 2004: 8th European Conference on Computer Vision, Prague, Czech Republic, May 11–14, 2004. Proceedings, Part I, volume 3021 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21984-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3021.htm;
 http://www.springerlink.
com/content/978-3-540-
21984-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3021;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b97865.

Pajdla:2004:CVEb

[PM04b] Tomás Pajdla and Jirà Matas, editors. Computer Vision — ECCV 2004: 8th European Conference on Computer Vision, Prague, Czech Republic, May 11–14, 2004. Proceedings, Part II, volume 3022 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21983-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3022.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21983-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3022;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97866.

Pajdla:2004:CVEc

[PM04c] Tomás Pajdla and Jirà Matas, editors. Computer Vision — ECCV 2004: 8th European Conference on Computer Vision, Prague, Czech Republic, May 11–14, 2004.

Proceedings, Part III, volume 3023 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21982-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3023.htm;
 http://www.springerlink.
com/content/978-3-540-
21982-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3023;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97871.

Pajdla:2004:CVEd

[PM04d] Tomás Pajdla and Jirà Matas, editors. Computer Vision — ECCV 2004: 8th European Conference on Computer Vision, Prague, Czech Republic, May 11–14, 2004. Proceedings, Part IV, volume 3024 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21981-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3024.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21981-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3024;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b97873.

Paech:2008:IRA

[PM08] Barbara Paech and Craig Martell, editors. Innovations for Requirement Analysis. From Stakeholders’ Needs to Formal Designs: 14th Monterey Workshop 2007, Monterey, CA, USA, September 10–13, 2007. Revised Selected Papers, volume 5320 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89777-1 (print), 3-540-89778-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89778-1.

Principe:2009:ASO

[PM09] José C. Príncipe and Risto Miikkulainen, editors. Advances in Self-Organizing Maps: 7th International Workshop, WSOM 2009, St. Augustine, FL, USA, June 8–10, 2009. Proceedings, volume 5629 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02396-7 (print), 3-642-02397-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02397-2.

Pelachaud:2007:IVA

[PMA^{+}07] Catherine Pelachaud, Jean-Claude Martin, Elisabeth André, Gérard Chollet, Kostas Karpouzis, and Danielle Pelé, editors. Intelligent Virtual Agents: 7th International Conference, IVA 2007 Paris, France, September 17–19, 2007 Proceedings, volume 4722 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74996-9 (print), 3-540-74997-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        74997-4.

Pointcheval:2006:CNS

[PMC06] David Pointcheval, Yi Mu, and Kefei Chen, editors. Cryptology and Network Security: 5th International Conference, CANS 2006, Suzhou, China, December 8–10, 2006. Proceedings, volume 4301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49462-6 (print), 3-540-49463-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49463-8.

Pichler:1990:CAS

[PMD90] Franz Pichler and Roberto Moreno-Diaz, editors. Computer aided systems theory — EUROCAST ’89: a selection of papers from the International Workshop EU-ROCAST ’89, Las Palmas, Spain, February 26–March 4, 1989: proceedings, volume 410 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52215-8 (USA).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .I585 1989. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0410.htm;
 http://www.springerlink.
com/content/978-0-387-
52215-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=410.

Pichler:1992:CAS

[PMD92] Franz Pichler and Roberto Moreno-Diaz, editors. Computer aided systems theory, EUROCAST ’91: a selection of papers from the Second International Workshop on Computer Aided Systems Theory, Krems, Austria, April 15–19, 1991: proceedings, volume 585 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 0-387-55354-1 (New York), 3-540-55354-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .I62 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0585.
htm;   http://www.springerlink.
com/content/978-0-387-
55354-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=585.

Pichler:1994:CAS

[PMD94] Franz Pichler and Roberto Moreno-Diaz, editors. Computer aided systems theory — EUROCAST ’93: a selection of papers from the Third International Workshop on Computer Aided Systems Theory, Las Palmas, Spain, February 22–26, 1993: proceedings, volume 763 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57601-0 (Berlin), 0-387-57601-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .I62 1993.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0763.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        57601-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=763.

Pichler:1997:CAS

[PMD97] Franz Pichler and Roberto Moreno-Diaz, editors. Computer aided systems theory EUROCAST ’97: a selection of papers from the 6th International Workshop on Computer Aided Systems Theory, Las Palmas de Gran Canaria, Spain, February 1997: proceedings, volume 1333 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63811-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345.I583 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1333.
htm;   http://www.springerlink.
com/content/978-3-540-
63811-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1333.

Pichler:1996:CAS

[PMDA96] Franz Pichler, Roberto Moreno-Diaz, and R. F. (Rudolf F.) Albrecht, editors. Computer aided systems theory — EU-ROCAST ’95: a selection of papers from the Fifth International Workshop on Computer Aided Systems Theory, Innsbruck, Austria, May 22–25, 1995: proceedings, volume 1030 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60748-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .I62 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1030.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60748-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1030.

Pichler:2000:CAS

[PMDK00] Franz Pichler, Roberto Moreno-Diaz, and Peter Kopacek, editors. Computer aided systems theory — EURO-CAST’99: a selection of papers from the 7th International Workshop on Computer Aided Systems Theory, Vienna, Austria, September 29 — October 2, 1999: Proceedings, volume 1798 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67822-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .I62 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1798.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67822-9;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1798.

Priami:2005:TCS

[PMGO05] Corrado Priami, Emanuela Merelli, Pablo Gonzalez, and Andrea Omicini, editors. Transactions on Computational Systems Biology III, volume 3737 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30883-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH506 .T732 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30883-6;   http://www.
springerlink.com/openurl.
asp?genre=book\%26isbn=
978-3-540-30883-6;
http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3737.

Paton:2009:DIL

[PMH09] Norman W. Paton, Paolo Missier, and Cornelia Hedeler, editors. Data Integration in the Life Sciences: 6th International Workshop, DILS 2009, Manchester, UK, July 20–22, 2009. Proceedings, volume 5647 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02878-0 (print), 3-642-02879-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02879-3.

Petta:2007:MST

[PMKG07] Paolo Petta, Jörg P. Müller, Matthias Klusch, and Michael Georgeff, editors. Multiagent System Technologies: 5th German Conference, MATES 2007, Leipzig, Germany, September 24– 26, 2007. Proceedings, volume 4687 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74948-9 (print), 3-540-74949-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74949-3.

Parsons:2006:AMA

[PMMR06] Simon Parsons, Nicolas Maudet, Pavlos Moraitis, and Iyad Rahwan, editors. Argumentation in Multi-Agent Systems: Second International Workshop, ArgMAS 2005 Utrecht, The Netherlands, July 26, 2005 Revised Selected and Invited Papers, volume 4049 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36355-6 (print), 3-540-36356-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36356-9.

Parr:2006:API

[PMÓ06] Gerard Parr, David Malone, and Mícheál Ó Foghlú, editors. Autonomic Principles of IP Operations and Management: 6th IEEE International Workshop on IP Operations and Management, IPOM 2006, Dublin, Ireland, October 23–25, 2006. Proceedings, volume 4268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47701-2 (print), 3-540-47702-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        47702-0.

Prevost:2008:ANN

[PMS08] Lionel Prevost, Simone Marinai, and Friedhelm Schwenker, editors. Artificial Neural Networks in Pattern Recognition: Third IAPR Workshop, ANNPR 2008 Paris, France, July 2–4, 2008 Proceedings, volume 5064 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69938-4 (print), 3-540-69939-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69939-2.

Pedersen:2009:DWK

[PMT09] Torben Bach Pedersen, Mukesh K. Mohania, and A. Min Tjoa, editors. Data Warehousing and Knowledge Discovery: 11th International Conference, DaWaK 2009 Linz, Austria, August 31–September 2, 2009 Proceedings, volume 5691 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03729-1 (print), 3-642-03730-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03730-6.

Pedersen:2010:DWK

[PMT10] Torben Bach Pedersen, Mukesh K. Mohania, and A. Min Tjoa, editors. Data Warehousing and Knowledge Discovery: 12th International Conference, DAWAK 2010, Bilbao, Spain, August/ September 2010. Proceedings, volume 6263 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15104-3 (print), 3-642-15105-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15105-7.

Pidduck:2002:AIS

[PMWO02] Anne Banks Pidduck, John Mylopoulos, Carson C. Woo, and M. Tamer Ozsu, editors. Advanced Information Systems Engineering: 14th International Conference, CAiSE 2002 Toronto, Canada, May 27–31, 2002. Proceedings, volume 2348 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43738-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .N67 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2348.htm;
 http://www.springerlink.
com/content/978-3-540-
43738-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2348.

Pfeiffer:1993:CST

[PN93] Heather D. Pfeiffer and Tim Nagle, editors. Conceptual structures: theory and implementation: 7th annual workshop, Las Cruces, NM, USA, July 8–10, 1992: proceedings, volume 754 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57454-9 (New York), 3-540-57454-9 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2 .C67 1993. DM66.00. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t0754.htm;
 http://www.springerlink.
com/content/978-0-387-
57454-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=754.

Based on papers from the Seventh Annual Workshop on Conceptual Graphs, held at New Mexico State University, Las Cruces, NM.

Pfaltz:2004:AGT

[PNB04] John L. Pfaltz, Manfred Nagl, and Boris Böhlen, editors. Applications of Graph Transformations with Industrial Relevance: Second International Workshop, AGTIVE 2003, Charlottesville, VA, USA, September 27–October 1, 2003, Revised Selected and Invited Papers, volume 3062 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22120-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3062.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        22120-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3062;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98116.

Petkova-Nikova:2012:PKI

[PNPP12] Svetla Petkova-Nikova, Andreas Pashalidis, and Günther Pernul, editors. Public Key Infrastructures, Services and Applications: 8th European Workshop, EuroPKI 2011, Leuven, Belgium, September 15– 16, 2011, Revised Selected Papers, volume 7163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29803-6 (print), 3-642-29804-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29804-2.

Podelski:1995:CPB

[Pod95] Andreas Podelski, editor. Constraint programming: basics and trends: 1994 Chatillon Spring School, Chatillonsur-Seine, France, May 16–20, 1994: Selected Papers, volume 910 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59155-9 (Berlin), 0-387-59155-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C67 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0910.htm;
 http://www.springerlink.
com/content/978-3-540-
59155-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=910.

Papers presented at the 22nd Spring School in Theoretical Computer Science in Chatillon-sur-Seine, France, May 16– 20, 1994.

Pointcheval:2006:TCC

[Poi06] David Pointcheval, editor. Topics in Cryptology

— CT-RSA 2006: The Cryptographers’ Track at the RSA Conference 2006, San Jose, CA, USA, February 13– 17, 2006, proceedings, volume 3860 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31033-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        31033-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3860.

Polak:1981:CSV

[Pol81] Wolfgang Polak. Compiler specification and verification, volume 124 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10886-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 269 pp. LCCN QA76.6 .P648. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0124.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        10886-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=124.

“A revised version of my Ph.D. thesis”– pref.

Poli:1986:AAA

[Pol86] Alain Poli, editor. Applied algebra, algorithmics and error-correcting codes: 2nd International Conference, AAECC-2, Toulouse, France, October 1–5, 1984: proceedings, volume 228 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16767-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A341 1984. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0228.htm;
 http://www.springerlink.
com/content/978-0-387-
16767-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=228.

Proceedings in English; one contribution in French.

Polychronopoulos:1997:HPC

[Pol97] Constantine D. Polychronopoulos, editor. High performance computing: International Symposium, ISHPC’97 Fukuoka, Japan, November 4–6, 1997: proceedings, volume 1336 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63766-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88.I585 1997. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1336.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63766-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1336.

Poli:1999:GPS

[Pol99a] Riccardo Poli, editor. Genetic programming: Second European Workshop, EuroGP ’99, Göteborg, Sweden, May 26–27, 1999: Proceedings, volume 1598 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65899-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.623 .E97 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1598.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65899-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1598.

Polychronopoulos:1999:HPC

[Pol99b] C. D. Polychronopoulos, editor. High performance computing: Second International Symposium, ISHPC’99, Kyoto, Japan, May 26– 28, 1999: Proceedings, volume 1615 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65969-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I585 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1615.htm;
 http://www.springerlink.
com/content/978-3-540-
65969-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1615.

Poli:2000:GPT

[Pol00] Riccardo Poli, editor.

Genetic programming: Third European Conference, EuroGP 2000, Edinburgh, Scotland, UK, April 15–16, 2000: Proceedings, volume 1802 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67339-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1802. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1802.htm;
 http://www.springerlink.
com/content/978-3-540-
67339-2;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1802.

Pollefeys:2001:SIS

[Pol01] Marc Pollefeys, editor. 3D structures from images — SMILE 2000: Second European Workshop on 3D Structure from Multiple Images of Large-Scale Environments, Dublin, Ireland, July 11–2, 2000: Revised Papers, volume 2018 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41845-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Electronic book. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2018.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41845-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2018.

Pomerance:1988:ACC

[Pom88] Carl Pomerance, editor. Advances in cryptology

— CRYPTO ’87: proceedings, volume 293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-18796-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C791 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0293.htm;
 http://www.springerlink.
com/content/978-0-387-
18796-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=293.

CRYPTO ’87, a Conference on the Theory and Applications of Cryptographic Techniques, held at the University of California, Santa Barbara … August 16–20, 1987.

Papadopouli:2009:TMA

[POP09] Maria Papadopouli, Philippe Owezarski, and Aiko Pras, editors. Traffic Monitoring and Analysis: First International Workshop, TMA 2009, Aachen, Germany, May 11, 2009. Proceedings, volume 5537 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01644-8 (print), 3-642-01645-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01645-5.

Porto:1998:EPV

[Por98] V. W. (V. William) Porto, editor. Evolutionary programming VII: 7th International Conference, EP98, San Diego, California, USA, March 25–27, 1998: Proceedings, volume 1447 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64891-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .C66 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1447.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64891-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1447.

Pieprzyk:2000:IST

[POS00] Josef Pieprzyk, Eiji Okamoto, and Jennifer Seberry, editors. Information Security: Third International Workshop, ISW 2000, Wollongong, Australia, December 20– 21, 2000. Proceedings, volume 1975 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41416-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1975. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1975.htm;
 http://www.springerlink.
com/content/978-3-540-
41416-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1975.

Perner:1999:MLD

[PP99] Petra Perner and Maria Petrou, editors. Machine learning and data mining in pattern recognition: First International Workshop, MLDM’99, Leipzig, Germany, September 16–18, 1999: Proceedings, volume 1715 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66599-4 (softcover), 3-540-66596-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1715. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1715.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66599-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1715.

Palanque:2001:ISD

[PP01] Philippe Palanque and Fabio Paternò, editors. Interactive systems: design, specification, and verification: 7th International Workshop, DSV-IS 2000, Limerick, Ireland, June 5– 6, 2000: Revised Papers, volume 1946 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41663-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.I58 D78 2000; QA267.A1 L43 no.1946. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1946.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41663-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1946.

Pinkston:2003:HPC

[PP03] Timothy Mark Pinkston and Viktor K. Prasanna, editors. High Performance Computing— HiPC 2003: 10th International Conference, Hyderabad, India, December 17–20, 2003: Proceedings, volume 2913 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20626-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I566 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2913.htm;
 http://www.springerlink.
com/content/978-3-540-
20626-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2913;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94479.

Puntonet:2004:ICA

[PP04] Carlos G. Puntonet and Alberto Prieto, editors. Independent Component Analysis and Blind Signal Separation: Fifth International Conference, ICA 2004, Granada, Spain, September 22– 24, 2004: Proceedings, volume 3195 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23056-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3195.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23056-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3195;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100528.

Priami:2006:TCSc

[PP06] Corrado Priami and Gordon Plotkin, editors. Transactions on Computational Systems Biology VI, volume 4220 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45779-8 (print), 3-540-46236-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46236-1.

Pogodalla:2011:LAC

[PP11] Sylvain Pogodalla and Jean-Philippe Prost, editors. Logical Aspects of Computational Linguistics: 6th International Conference, LACL 2011, Montpellier, France, June 29 — July 1, 2011. Proceedings, volume 6736 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22220-X (print), 3-642-22221-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22221-4.

Pankratius:2012:MSE

[PP12a] Victor Pankratius and Michael Philippsen, editors. Multicore Software Engineering, Performance, and Tools: International Conference, MSEPT 2012, Prague, Czech Republic, May 31 — June 1, 2012. Proceedings, volume 7303 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31201-2 (print), 3-642-31202-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31202-1.

Pena:2012:TFP

[PP12b] Ricardo Peña and Rex Page, editors. Trends in Functional Programming: 12th International Symposium, TFP 2011, Madrid, Spain, May 16–18, 2011, Revised Selected Papers, volume 7193 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32036-8 (print), 3-642-32037-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32037-8.

Pinz:2012:PRJ

[PPBL12] Axel Pinz, Thomas Pock, Horst Bischof, and Franz Leberl, editors. Pattern Recognition: Joint 34th DAGM and 36th OAGM Symposium, Graz, Austria, August 28–31, 2012. Proceedings, volume 7476 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32716-8 (print), 3-642-32717-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32717-9.

Palmirani:2012:AAC

[PPCS12] Monica Palmirani, Ugo Pagallo, Pompeu Casanovas, and Giovanni Sartor, editors. AI Approaches to the Complexity of Legal Systems. Models and Ethical Challenges for Legal Systems, Legal Language and Legal Ontologies, Argumentation and Software Agents: International Workshop AICOL-III, Held as Part of the 25th IVR Congress, Frankfurt am Main, Germany, August 15–16, 2011. Revised Selected Papers, volume 7639 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35730-X (print), 3-642-35731-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35731-2.

Priami:2012:TCS

[PPdV12] Corrado Priami, Ion Petre, and Erik de Vink, editors. Transactions on Computational Systems Biology XIV: Special Issue on Computational Models for Cell Processes, volume 7625 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35523-4 (print), 3-642-35524-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35524-0.

Pujolle:2000:NBC

[PPF^{+}00] Guy Pujolle, Harry Perros, Serge Fdida, Ulf Körner, and Ioannis Stavrakakis, editors. Networking 2000: broadband communications, high performance networking, and performance of communication networks: IFIP-TC6/European Commission International Conference, Paris, France, May 14–19, 2000: Proceedings, volume 1815 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67506-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.4 .I33 2000. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t1815.htm;
http://www.springerlink.
com/content/978-3-540-
67506-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1815.

Priss:2007:CSK

[PPH07] Uta Priss, Simon Polovina, and Richard Hill, editors. Conceptual Structures: Knowledge Architectures for Smart Applications: 15th International Conference on Conceptual Structures, ICCS 2007, Sheffield, UK, July 22–27, 2007. Proceedings, volume 4604 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73680-8 (print), 3-540-73681-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73681-3.

Paun:2010:MCI

[PPJRN^{+}10] Gheorghe Păun, Mario J.

Pérez-Jiménez, Agustín Riscos-Núñez, Grzegorz Rozenberg, and Arto Salomaa, editors. Membrane Computing: 10th International Workshop, WMC 2009, Curtea de Arges, Romania, August 24–27, 2009. Revised Selected and Invited Papers, volume 5957 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11466-0 (print), 3-642-11467-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11467-0.

Prince:2009:IPM

[PPM09] Jerry L. Prince, Dzung L. Pham, and Kyle J. Myers, editors. Information Processing in Medical Imaging: 21st International Conference, IPMI 2009, Williamsburg, VA, USA, July 5–10, 2009. Proceedings, volume 5636 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02497-1 (print), 3-642-02498-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02498-6.

Paiva:2007:ACI

[PPP07] Ana C. R. Paiva, Rui Prada, and Rosalind W. Picard, editors. Affective Computing and Intelligent Interaction: Second International Conference, ACII 2007 Lisbon, Portugal, September 12–14, 2007 Proceedings, volume 4738 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74888-1 (print), 3-540-74889-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74889-2.

Papazoglou:2010:SRC

[PPPM10] Mike P. Papazoglou, Klaus Pohl, Michael Parkin, and Andreas Metzger, editors. Service Research Challenges and Solutions for the Future Internet: S-Cube — Towards Engineering, Managing and Adapting Service-Based Systems, volume 6500 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17598-8 (print), 3-642-17599-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17599-2.

Pitt:1987:CTC

[PPR87] David Pitt, A. Poigne, and D. E. Rydeheard, editors. Category theory and computer science: Edinburgh, UK, September 7–9, 1987: proceedings, volume 283 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18508-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .C3881 1987. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0283.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        18508-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=283.

Petreschi:2003:ACI

[PPS03] Rossella Petreschi, Giuseppe Persiano, and Riccardo Silvestri, editors. Algorithms and Complexity: 5th Italian Conference, CIAC 2003, Rome, Italy, May 28–30, 2003: Proceedings, volume 2653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40176-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .A42 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2653.htm;
 http://www.springerlink.
com/content/978-3-540-
40176-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2653.

Also available via the World Wide Web.

Perner:1996:ASS

[PpWR96] Petra Perner, Patrick S-P. (Patrick Shen pei) Wang, and Azriel Rosenfeld, editors. Advances in structural and syntactical pattern recognition: 6th International Workshop, SSPR ’96, Leipzig, Germany, August 20–23, 1996: proceedings, volume 1121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61577-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 A28 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1121.

                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61577-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1121.

Proceedings of the 6th International Workshop on Structural and Syntactical Pattern Recognition.

Priami:2005:GCI

[PQ05] Corrado Priami and Paola Quaglia, editors. Global Computing: IST/FET International Workshop, GC 2004 Rovereto, Italy, March 9–12, 2004 Revised Selected Papers, volume 3267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24101-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24101-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3267;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b103251.

Pogodalla:2011:LGE

[PQR11] Sylvain Pogodalla, Myriam Quatrini, and Christian Retoré, editors. Logic and Grammar: Essays Dedicated to Alain Lecomte on the Occasion of His 60th Birthday, volume 6700 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21489-4 (print), 3-642-21490-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21490-5.

Paul:1984:ISP

[PR84] Manfred Paul and B. Robinet, editors. International Symposium on Programming: proceedings, 6th colloquium, Toulouse, April 17–19, 1984, volume 167 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CODEN LNCSD9. ISBN 3-540-12925-1, 0-387-12925-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6.I575 1984. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0167.
htm;   http://www.springerlink.
com/content/978-3-540-

                                        12925-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=167.

Pardalos:1987:CGO

[PR87] P. M. (Panos M.) Pardalos and J. B. Rosen. Constrained global optimization: algorithms and applications, volume 268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-18095-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 143 pp. LCCN QA402.5 .P3831 1987. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0268.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        18095-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=268.

Puech:1996:SAS

[PR96] Claude Puech and Rudiger Reischuk, editors. STACS 96: 13th Annual Symposium on Theoretical Aspects of Computer Science, Grenoble, France, February 22–24, 1996: proceedings, volume 1046 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60922-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S958 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1046.htm;
 http://www.springerlink.
com/content/978-3-540-
60922-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1046.

Privara:1997:MFC

[PR97] Igor Privara and Peter Ružička, editors. Mathematical foundations of computer science 1997: 22nd International Symposium, MFCS ’97, Bratislava, Slovakia, August 25–29, 1997: proceedings, volume 1295 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63437-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 M387 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1295.
htm;   http://www.springerlink.
com/content/978-3-540-
63437-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1295.

Preneel:1998:SAA

[PR98a] Bart Preneel and Vincent Rijmen, editors. State of the Art in Applied Cryptography: Course on Computer Security and Industrial Cryptography, Leuven, Belgium, June 3–6, 1997. Revised Lectures, volume 1528 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65474-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25S735 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1528.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65474-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1528.

Pritchard:1998:EPP

[PR98b] David Pritchard and Jeff Reeve, editors. Euro-Par’98 parallel processing: 4th International Euro-Par Conference, Southampton, UK, September 1–4, 1998: Proceedings, volume 1470 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64952-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5535 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1470.htm;
 http://www.springerlink.
com/content/978-3-540-
64952-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1470.

Porto:2000:CLM

[PR00] A. Porto and Gruia-Catalin Roman, editors. Coordination Languages and Models: Fourth International Conference, COORDINATION 2000, Lymasol, Cyprus, September 11–13, 2000. Proceedings, volume 1906 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41020-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .C695 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1906.htm;
 http://www.springerlink.
com/content/978-3-540-
41020-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1906.

Pacholski:2001:STP

[PR01] Leszek Pacholski and Peter Ružička, editors. SOFSEM 2001: Theory and Practice of Informatics: 28th Conference on Current Trends in Theory and Practice of Informatics Piešt’any, Slovak Republic, November 24– December 1, 2001. Proceedings, volume 2234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42912-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .S62 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2234.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42912-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2234.

Pandya:2003:FTF

[PR03a] Paritosh K. Pandya and Jaikumar Radhakrishnan, editors. FST TCS 2003: Foundations of Software Technology and Theoretical Computer Science, volume 2914 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20680-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2914.htm;
 http://www.springerlink.
com/content/978-3-540-
20680-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2914;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94618.

Perner:2003:MLD

[PR03b] Petra Perner and Azriel Rosenfeld, editors. Machine learning and data mining in pattern recognition: Third International Conference, MLDM 2003, Leipzig, Germany, July 25 5–7, 2003: Proceedings, volume 2734 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40504-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q327 .I67 2003 Stacks. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2734.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40504-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2734.

Pelc:2005:SIC

[PR05] Andrzej Pelc and Michel Raynal, editors. Structural Information and Communication Complexity: 12th International Colloquium, SIROCCO 2005 Mont Saint-Michel, France, May 24–26, 2005 Proceedings, volume 3499 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26052-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26052-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3499;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136628.

Paletta:2007:ACS

[PR07] Lucas Paletta and Erich Rome, editors. Attention in Cognitive Systems. Theories and Systems from an Interdisciplinary Viewpoint: 4th International Workshop on Attention in Cognitive Systems, WAPCV 2007 Hyderabad, India, January 8, 2007 Revised Selected Papers, volume 4840 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77342-8 (print), 3-540-77343-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77343-6.

Paech:2008:REF

[PR08] Barbara Paech and Colette Rolland, editors. Requirements Engineering: Foundation for Software Quality: 14th International Working Conference, REFSQ 2008 Montpellier, France, June 16–17, 2008 Proceedings, volume 5025 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69060-3 (print), 3-540-69062-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69062-7.

Pardalos:2011:EAI

[PR11] Panos M. Pardalos and Steffen Rebennack, editors. Experimental Algorithms: 10th International Symposium, SEA 2011, Kolimpari, Chania, Crete, Greece, May 5–7, 2011. Proceedings, volume 6630 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20661-1 (print), 3-642-20662-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20662-7.

Pipek:2009:EUD

[PRdRW09] Volkmar Pipek, Mary Beth Rosson, Boris de Ruyter, and Volker Wulf, editors. End-User Development: 2nd International Symposium, IS-EUD 2009, Siegen, Germany, March 2–4, 2009. Proceedings, volume 5435 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00425-3 (print), 3-642-00427-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00427-8.

Preneel:1995:FSE

[Pre95] Bart Preneel, editor. Fast software encryption: second international workshop, Leuven, Belgium, December 14–16, 1994: proceedings, volume 1008 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60590-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 F37 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1008.htm;
http://www.springerlink.
com/content/978-3-540-
60590-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1008.

Preneel:2000:ACE

[Pre00] Bart Preneel, editor.

Advances in cryptology: EUROCRYPT 2000: International Conference on the Theory and Application of Cryptographic Techniques, Bruges, Belgium, May 14–18, 2000: Proceedings, volume 1807 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67517-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1807.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1807.
htm;   http://www.springerlink.
com/content/978-3-540-
67517-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1807.

Preneel:2002:TCC

[Pre02] Bart Preneel, editor. Topics in cryptology, CT-RSA 2002: the Cryptographers’ Track at RSA Conference 2002, San Jose, CA, USA, February 18–22, 2002: Proceedings, volume 2271 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43224-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 R753 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2271.
htm;   http://www.springerlink.
com/content/978-3-540-
43224-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2271.

Preneel:2009:PCA

[Pre09] Bart Preneel, editor.

Progress in Cryptology — AFRICACRYPT 2009: Second International Conference on Cryptology in Africa, Gammarth, Tunisia, June 21–25, 2009. Proceedings, volume 5580 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02383-5 (print), 3-642-02384-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02384-2.

Pizzuti:2009:ECM

[PRG09] Clara Pizzuti, Marylyn D. Ritchie, and Mario Giacobini, editors. Evolutionary Computation, Machine Learning and Data Mining in Bioinformatics: 7th European Conference, Evo-BIO 2009 Tübingen, Germany, April 15– 17, 2009 Proceedings, volume 5483 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01183-7 (print), 3-642-01184-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01184-9.

Pizzuti:2010:ECM

[PRG10] Clara Pizzuti, Marylyn D. Ritchie, and Mario Giacobini, editors. Evolutionary Computation, Machine Learning and Data Mining in Bioinformatics: 8th European Conference, EvoBIO 2010, Istanbul, Turkey, April 7–9, 2010. Proceedings, volume 6023 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12210-8 (print), 3-642-12211-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12211-8.

Pizzuti:2011:ECM

[PRG11] Clara Pizzuti, Marylyn D. Ritchie, and Mario Giacobini, editors. Evolutionary Computation, Machine Learning and Data Mining in Bioinformatics: 9th European Conference, Evo-BIO 2011, Torino, Italy, April 27–29, 2011. Proceedings, volume 6623 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20388-4 (print), 3-642-20389-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20389-3.

Petriu:2010:MDEa

[PRH10a] Dorina C. Petriu, Nicolas Rouquette, and Øystein Haugen, editors. Model Driven Engineering Languages and Systems: 13th International Conference, MODELS 2010, Oslo, Norway, October 3–8, 2010, Proceedings, Part I, volume 6394 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16144-8 (print), 3-642-16145-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16145-2.

Petriu:2010:MDEb

[PRH10b] Dorina C. Petriu, Nicolas Rouquette, and Øystein Haugen, editors. Model Driven Engineering Languages and Systems: 13th International Conference, MODELS 2010, Oslo, Norway, October 3–8, 2010, Proceedings, Part II, volume 6395 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16128-6 (print), 3-642-16129-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16129-2.

Prieto:1991:ANN

[Pri91] A. Prieto, editor. Artificial neural networks: International Workshop, IWANN ’91, Granada, Spain, September 17– 19, 1991: proceedings, volume 540 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54537-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I586 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0540.htm;
 http://www.springerlink.

com/content/978-0-387-
54537-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=540.

Priami:2003:CMS

[Pri03a] Corrado Priami, editor. Computational methods in systems biology: First International Workshop, CMSB 2003, Rovereto, Italy, February 24–26, 2003: Proceedings, volume 2602 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00605-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH506 .I525 2003; QH506 .I613 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2602.
htm;   http://www.springerlink.
com/content/978-3-540-
00605-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2602.

Also available via the World Wide Web.

Priami:2003:GCP

[Pri03b] Corrado Priami, editor.

Global Computing: Programming Environments, Languages, Security, and Analysis of Systems: IST/FET International Workshop, GC 2003, Rovereto, Italy, February 9–14, 2003: Revised Papers, volume 2874 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20583-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 G373 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2874.htm;
 http://www.springerlink.
com/content/978-3-540-
20583-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2874;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94264.

Priami:2005:TCSa

[Pri05] Corrado Priami, editor. Transactions on Computational Systems Biology I, volume 3380 of Lecture Notes in Bioinformatics and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25422-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 111 pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25422-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3380;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107357.

Priami:2006:CMS

[Pri06] Corrado Priami, editor. Computational Methods in Systems Biology: International Conference, CMSB 2006, Trento, Italy, October 18–19, 2006. Proceedings, volume 4210 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46166-3 (print), 3-540-46167-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46167-8.

Priami:2007:TCS

[Pri07] Corrado Priami. Transactions on Computational Systems Biology VIII, volume 4780 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-76639-1. 101 pp. LCCN QH506 .T736 2007eb. URL http:

//www.springerlink.
com/openurl.asp?genre=
book\%26isbn=978-3-
540-76638-4.

Priami:2008:TCSa

[Pri08] Corrado Priami, editor. Transactions on Computational Systems Biology IX, volume 5121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88764-4 (print), 3-540-88765-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88765-2.

Pitt:1995:CTC

[PRJ95] David (David H.) Pitt, D. E. (David E.) Rydeheard, and P. T. Johnstone, editors. Category theory and computer science: 6th International Conference, CTCS ’95, Cambridge, United Kingdom, August 7–11, 1995: proceedings, volume 953 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60164-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6.B53 1995. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0953.
htm;   http://www.springerlink.
com/content/978-3-540-
60164-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=953.

Peffers:2012:DSR

[PRK12] Ken Peffers, Marcus Rothenberger, and Bill Kuechler, editors. Design Science Research in Information Systems. Advances in Theory and Practice: 7th International Conference, DESRIST 2012, Las Vegas, NV, USA, May 14–15, 2012. Proceedings, volume 7286 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29862-1 (print), 3-642-29863-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29863-9.

Proietti:1996:LPS

[Pro96] Maurizio Proietti, editor. Logic program synthesis and transformation: 5th International Workshop, LOP-STR’95, Utrecht, The Netherlands, September 20–22, 1995: proceedings, volume 1048 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60939-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I57 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1048.htm;
 http://www.springerlink.
com/content/978-3-540-
60939-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1048.

Prouff:2011:SCR

[Pro11] Emmanuel Prouff, editor. Smart Card Research and Advanced Applications: 10th IFIP WG 8.8/11.2 International Conference, CARDIS 2011, Leuven, Belgium, September 14–16, 2011, Revised Selected Papers, volume 7079 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-27256-8 (print), 3-642-27257-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27257-8.

Privara:1994:MFC

[PRR94] Igor Privara, B. Rovan, and Peter Ružička, editors. Mathematical foundations of computer science 1994: 19th International Symposium, MFCS’94, Kosice, Slovakia, August 22–26, 1994: proceedings, volume 841 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58338-6 (Berlin), 0-387-58338-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 I54 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0841.htm;
http://www.springerlink.
com/content/978-3-540-
58338-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=841.

Prinz:2005:SMD

[PRR05] Andreas Prinz, Rick Reed, and Jeanne Reed, editors. SDL 2005:

Model Driven Systems Design: 12th International SDL Forum, Grimstad, Norway, June 20–23, 2005. Proceedings, volume 3530 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26612-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.S35 S35 2005. URL http:

//www.springerlink.
com/content/978-3-540-
26612-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3530.

Penkler:2006:SAT

[PRT06] Dave Penkler, Manfred Reitenspiess, and Francis Tam, editors. Service Availability: Third International Service Availability Symposium, ISAS 2006, Helsinki, Finland, May 15–16, 2006. Revised Selected Papers, volume 4328 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68724-6 (print), 3-540-68725-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68725-2.

Pahl:2011:NOI

[PRV11] Julia Pahl, Torsten Reiners, and Stefan Voß, editors. Network Optimization: 5th International Conference, INOC 2011, Hamburg, Germany, June 13–16, 2011. Proceedings, volume 6701 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21526-2 (print), 3-642-21527-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21527-8.

Pieprzyk:1993:DHA

[PS93] Josef Pieprzyk and Babak Sadeghiyan. Design of hashing algorithms, volume 756 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57500-6 (Berlin), 0-387-57500-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 194 pp. LCCN QA76.9.H36 P53 1993.

URL http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t0756.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        57500-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=756.

Penczek:1996:MFC

[PS96a] Wojciech Penczek and Andrzej Szalas, editors. Mathematical foundations of computer science 1996: 21st International Symposium, MFCS ’96, Crakow, Poland, September 2– 6, 1996: proceedings, volume 1113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61550-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 S96 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1113.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61550-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1113.

Pieprzyk:1996:ISP

[PS96b] Josef Pieprzyk and Jennifer Seberry, editors. Information security and privacy: first Australian conference, ACISP ’96, Wollongong, NSW, Australia, June 24–26, 1996: proceedings, volume 1172 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61991-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25A278 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1172.htm;
 http://www.springerlink.
com/content/978-3-540-
61991-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1172.

Panda:1997:CAS

[PS97a] Dhabaleswar K. Panda and Craig B. Stunkel, editors. Communication and architectural support for network-based parallel computing: First International Workshop, CANPC’97, San Antonio, Texas, USA, February 1997: proceedings, volume 1199 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62573-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.658 .C36 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1199.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62573-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1199.

Paun:1997:NTF

[PS97b] Gheorghe Păun and Arto Salomaa. New trends in formal languages: control, cooperation, and combinatorics, volume 1218 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62844-4. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 464 pp. LCCN QA76.7.N49 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1218.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62844-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1218.

Panda:1998:NBP

[PS98a] Dhabaleswar K. Panda and Craig B. Stunkel, editors. Network-based parallel computing: communication, architecture, and applications: second international workshop, CANPC ’98, Las Vegas, Nevada, USA, January 31–February 1, 1998, proceedings, volume 1362 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64140-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.C37 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1362.
htm;   http://www.springerlink.
com/content/978-3-540-
64140-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1362.

Polkowski:1998:RSC

[PS98b] Lech Polkowski and Andrzej Skowron, editors. Rough sets and current trends in computing: First International Conference, RSCTC ’98, Warsaw, Poland, June 22–26, 1998: Proceedings, volume 1424 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64655-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1424. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1424.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64655-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1424.

Pal:2002:ASC

[PS02a] Nikhil R. Pal and Michio Sugeno, editors. Advances in soft computing: AFSS 2002: 2002 AFSS International Conference on Fuzzy Systems, Calcutta, India, February 3–6, 2002: Proceedings, volume 2275 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43150-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S63 A3 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2275.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        43150-3;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2275.

Penttonen:2002:ATS

[PS02b] Martti Penttonen and E. Meineche Schmidt, editors. Algorithm theory — SWAT 2002: 8th Scandinavian Workshop on Algorithm Theory, Turku, Finland, July 3–5, 2002: Proceedings, volume 2368 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43866-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .S28 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2368.htm;
http://www.springerlink.
com/content/978-3-540-
43866-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2368.

Pfenning:2003:GPC

[PS03] Frank Pfenning and Yannis Smaragdakis, editors. Generative Programming and Component Engineering: Second International Conference, GPCE 2003, Erfurt, Germany, September 22–25, 2003: Proceedings, volume 2830 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20102-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.624 .G64 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2830.htm;
 http://www.springerlink.
com/content/978-3-540-
20102-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2830;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13639.

Paliouras:2004:GIA

[PS04a] Georgois Paliouras and Yasubumi Sakakibara, editors. Grammatical inference: algorithms and applications: 7th international colloquium, ICGI 2004, Athens, Greece, October 11–13, 2004: proceedings, volume 3264 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23410-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA267.3 .I55 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23410-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3264;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101520.

Persson:2004:AIS

[PS04b] Anne Persson and Janis Stirna, editors. Advanced Information Systems Engineering: 16th International Conference, CAiSE 2004, Riga, Latvia, June 7–11, 2004: Proceedings, volume 3084 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22151-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .C33 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3084.htm;
 http://www.springerlink.
com/content/978-3-540-
22151-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3084;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98058.

Peters:2005:TRSb

[PS05a] James F. Peters and Andrzej Skowron, editors. Transactions on rough sets III, volume 3400 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25998-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 461 pp. LCCN QA248 .T65 2005eb. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25998-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3400;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136502.

Peters:2005:TRS

[PS05b] James F. Peters and Andrzej Skowron, editors. Transactions on Rough Sets IV, volume 3700 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29830-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA248 .T73 2005b; Q335 .T72 2005b. URL http:

                                        //www.springerlink.

com/content/978-3-540-
29830-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3700.

Peters:2006:TRS

[PS06] James F. Peters and Andrzej Skowron, editors. Transactions on Rough Sets V, volume 4100 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39382-X (print), 3-540-39383-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
39383-2.

Perner:2007:AMD

[PS07a] Petra Perner and Ovidio Salvetti, editors. Advances in Mass Data Analysis of Signals and Images in Medicine, Biotechnology and Chemistry: International Conferences MDA 2006/ 2007, Leipzig, Germany, July 18, 2007. Selected Papers, volume 4826 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76299-X (print), 3-540-76300-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76300-0.

Prade:2007:SUM

[PS07b] Henri Prade and V. S. Subrahmanian, editors. Scalable Uncertainty Management: First International Conference, SUM 2007, Washington,DC, USA, October 10–12, 2007. Proceedings, volume 4772 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75407-5 (print), 3-540-75410-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75410-7.

Perner:2008:AMD

[PS08a] Petra Perner and Ovidio Salvetti, editors. Advances in Mass Data Analysis of Images and Signals in Medicine, Biotechnology, Chemistry and Food Industry: Third International Conference, MDA 2008 Leipzig, Germany, July 14, 2008 Proceedings, volume 5108 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70714-X (print), 3-540-70715-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70715-8.

Peters:2008:TRSa

[PS08b] James F. Peters and Andrzej Skowron, editors. Transactions on Rough Sets VIII, volume 5084 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85063-5 (print), 3-540-85064-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=5084.

Palsberg:2009:SAI

[PS09a] Jens Palsberg and Zhendong Su, editors. Static Analysis: 16th International Symposium, SAS 2009, Los Angeles, CA, USA, August 9–11, 2009. Proceedings, volume 5673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03236-2 (print), 3-642-03237-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03237-0.

Polleres:2009:WRR

[PS09b] Axel Polleres and Terrance Swift, editors. Web Reasoning and Rule Systems: Third International Conference, RR 2009, Chantilly, VA, USA, October 25–26, 2009, Proceedings, volume 5837 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05081-6 (print), 3-642-05082-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05082-4.

Purvis:2009:CMS

[PS09c] Maryam Purvis and Bastin Tony Roy Savarimuthu, editors. Computer-Mediated Social Networking: First International Conference, ICCMSN 2008, Dunedin, New Zealand, June 11–13, 2008, Revised Selected Papers, volume 5322 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02275-8 (print), 3-642-02276-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02276-0.

Peters:2010:TRSa

[PS10] James F. Peters and Andrzej Skowron, editors. Transactions on Rough Sets XI, volume 5946 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11478-4 (print), 3-642-11479-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11479-3.

Pirrone:2011:AAI

[PS11a] Roberto Pirrone and Filippo Sorbello, editors. AI*IA 2011: Artificial Intelligence Around Man and Beyond: XIIth International Conference of the Italian Association for Artificial Intelligence, Palermo, Italy, September 15–17, 2011. Proceedings, volume 6934 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23953-6 (print), 3-642-23954-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23954-0.

Przytycka:2011:ABI

[PS11b] Teresa M. Przytycka and Marie-France Sagot, editors. Algorithms in Bioinformatics: 11th International Workshop, WABI 2011, Saarbrücken, Germany, September 5– 7, 2011. Proceedings, volume 6833 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23037-7 (print), 3-642-23038-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23038-7.

Pattinson:2012:CMC

[PS12a] Dirk Pattinson and Lutz Schröder, editors. Coalgebraic Methods in Computer Science: 11th International Workshop, CMCS 2012, Colocated with ETAPS 2012, Tallinn, Estonia, March 31 — April 1, 2012, Revised Selected Papers, volume 7399 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32783-4 (print), 3-642-32784-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32784-1.

Peters:2012:TRS

[PS12b] James F. Peters and Andrzej Skowron, editors. Transactions on Rough Sets XV, volume 7255 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31902-5 (print), 3-642-31903-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31903-7.

Prouff:2012:CHE

[PS12c] Emmanuel Prouff and Patrick Schaumont, editors. Cryptographic Hardware and Embedded Systems — CHES 2012: 14th International Workshop, Leuven, Belgium, September 9–12, 2012. Proceedings, volume 7428 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33026-6 (print), 3-642-33027-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33027-8.

Peters:2011:TRSa

[PSC^{+}11] James F. Peters, Andrzej Skowron, Chien-Chung Chan, Jerzy W. Grzymala-Busse, and Wojciech P. Ziarko, editors. Transactions on Rough Sets XIII, volume 6499 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18301-8 (print), 3-642-18302-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18302-7.

Peters:2005:TRSa

[PSD^{+}05] James F. Peters, Andrzej Skowron, Didier Dubois, Jerzy W. Grzymała-Busse, Masahiro Inuiguchi, and Lech Polkowski. Transactions on rough sets II: rough sets and fuzzy sets, volume 3135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-23990-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 361 pp. LCCN QA248.5 .T735 2004.

URL http://www.springerlink.
com/content/978-3-540-
23990-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3135.

Peters:2007:TRSa

[PSD^{+}07] James F. Peters, Andrzej Skowron, Ivo Düntsch, Jerzy Grzymała-Busse, Ewa Orłowska, and Lech Polkowski, editors. Transactions on rough sets VI: commemorating the life and work of Zdzisław Pawlak, Part 1, volume 4374 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-71198-8. xii + 498 pp. LCCN QA248 .T736 2007. URL http:

                                        //link.springer.com/
                                        book/10.1007/978-3-
                                        540-71200-8/;   http:
                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        book\%26id=doi:10.1007/
                                        978-3-540-71200-8.

Pescape:2012:TMA

[PSD12] Antonio Pescapè, Luca Salgarelli, and Xenofontas Dimitropoulos, editors. Traffic Monitoring and Analysis: 4th International Workshop, TMA 2012, Vienna, Austria, March 12, 2012. Proceedings, volume 7189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28533-3 (print), 3-642-28534-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28534-9.

Panigrahi:2011:SEMa

[PSDS11a] Bijaya Ketan Panigrahi, Ponnuthurai Nagaratnam Suganthan, Swagatam Das, and Suresh Chandra Satapathy, editors. Swarm, Evolutionary, and Memetic Computing: Second International Conference, SEM-CCO 2011, Visakhapatnam, Andhra Pradesh, India, December 19–21, 2011, Proceedings, Part I, volume 7076 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-27171-5 (print), 3-642-27172-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27172-4.

Panigrahi:2011:SEMb

[PSDS11b] Bijaya Ketan Panigrahi, Ponnuthurai Nagaratnam Suganthan, Swagatam Das, and Suresh Chandra Satapathy, editors. Swarm, Evolutionary, and Memetic Computing: Second International Conference, SEM-CCO 2011, Visakhapatnam, Andhra Pradesh, India, December 19–21, 2011, Proceedings, Part II, volume 7077 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-27241-X (print), 3-642-27242-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27242-4.

Patt-Shamir:2010:SIC

[PSE10] Boaz Patt-Shamir and Tınaz Ekim, editors. Structural Information and Communication Complexity: 17th International Colloquium, SIROCCO 2010, Şirince, Turkey, June 7–11, 2010. Proceedings, volume 6058 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13283-9 (print), 3-642-13284-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13284-1.

Peters:2004:TRS

[PSGB^{+}04] James F. Peters, Andrzej Skowron, Jerzy W. Grzymała-Busse, Bożena Kostek, Roman W. Świniarski, and Marcin S. Szczuka, editors. Transactions on Rough Sets I, volume 3100 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22374-6. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 404 pp. LCCN Q335.T73 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3100.htm;
 http://www.springerlink.
com/content/978-3-540-
22374-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3100;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/

b98175. First volume of the new journal Transactions on Rough Sets.

Peters:2007:TRSb

[PSM^{+}07] James F. Peters, Andrzej Skowron, Victor W. Marek, Ewa Orłowska, Roman Słowinski, and Wojciech Ziarko, editors. Transactions on rough sets VII: commemorating the life and work of Zdzisław Pawlak, Part II, volume 4400 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CO-DEN LNCSD9. ISBN 3-540-71662-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer.com/
book/10.1007/978-3-

                                        540-71663-1;   http:
                                        //springerlink.metapress.
                                        com/openurl.asp?genre=
                                        issue\%26issn=0302-
                                        9743\%26volume=4400;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=4400.

Petrenko:2010:TSS

[PSM10] Alexandre Petrenko, Adenilso Simão, and José Carlos Maldonado, editors. Testing Software and Systems: 22nd IFIP WG 6.1 International Conference, ICTSS 2010, Natal, Brazil, November 8–10, 2010. Proceedings, volume 6435 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16572-9 (print), 3-642-16573-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16573-3.

Pieprzyk:1995:ACA

[PSN95] Josef Pieprzyk and Reihanah Safavi-Naini, editors. Advances in cryptology, ASIACRYPT ’94: 4th International Conference on the Theory and Application of Cryptology, Wollongong, Australia, November 28– December 1, 1994: proceedings, volume 917 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59339-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I555 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0917.
htm;   http://www.springerlink.
com/content/978-3-540-
59339-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=917.

Pieprzyk:1999:ISP

[PSNS99] Josef Pieprzyk, Reihanah Safavi-Naini, and Jennifer Seberry, editors. Information security and privacy: 4th Australasian Conference, ACISP ’99, Wollongong, NSW, Australia, April 7–9, 1999: Proceedings, volume 1587 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65756-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25A279 1999. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1587.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65756-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1587.

Persiano:2005:AOA

[PSO05] Giuseppe Persiano and Roberto Solis-Oba, editors. Approximation and online algorithms: second international workshop, WAOA 2004, Bergen, Norway, September 14–16, 2004: Revised selected papers, volume 3351 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24574-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W42 2004.

                                        URL http://springerlink.
                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3351;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        24574-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3351;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106130.

Padget:2002:AME

[PSP^{+}02] Julian A. Padget, Onn Shehory, David Parkes, Norman Sadeh, and William E. Walsh, editors. Agent-mediated electronic commerce IV: designing mechanisms and systems: AAMAS 2002 Workshop on Agent-Mediated Electronic Commerce, Bologna, Italy, July 16, 2002: Revised Papers, volume 2531 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00327-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.32 .I585 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2531.htm;
http://www.springerlink.
com/content/978-3-540-
00327-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2531.

Also available via the World Wide Web.

Patel-Schneider:2010:SWIa

[PSPH^{+}10a] Peter F. Patel-Schneider, Yue Pan, Pascal Hitzler, Peter Mika, Lei Zhang, Jeff Z. Pan, Ian Horrocks, and Birte Glimm, editors. The Semantic Web — ISWC 2010: 9th International Semantic Web Conference, ISWC 2010, Shanghai, China, November 7–11, 2010, Revised Selected Papers, Part I, volume 6496 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17745-X (print), 3-642-17746-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17746-0.

Patel-Schneider:2010:SWIb

[PSPH^{+}10b] Peter F. Patel-Schneider, Yue Pan, Pascal Hitzler, Peter Mika, Lei Zhang, Jeff Z. Pan, Ian Horrocks, and Birte Glimm, editors. The Semantic Web — ISWC 2010: 9th International Semantic Web Conference, ISWC 2010, Shanghai, China, November 7–11, 2010, Revised Selected Papers, Part II, volume 6497 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17748-4 (print), 3-642-17749-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17749-1.

Peters:2008:TRSb

[PSR08] James F. Peters, Andrzej Skowron, and Henryk Rybiński, editors. Transactions on Rough Sets IX, volume 5390 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89875-1 (print), 3-540-89876-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89876-4.

Puigjaner:1998:CPE

[PSS98] Ramon Puigjaner, Nunzio N. Savino, and Bartomeu Serra, editors. Computer performance evaluation: modelling techniques and tools: 10th International Conference, Palma de Mallorca, Spain, September 14–18, 1998: Proceedings, volume 1469 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64949-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.E94I55 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1469.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64949-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1469.

Parsons:2010:CMI

[PSS^{+}10a] Jeffrey Parsons, Motoshi Saeki, Peretz Shoval, Carson Woo, and Yair Wand, editors. Conceptual Modeling — ER 2010: 29th International Conference on Conceptual Modeling, Vancouver, BC, Canada, November 1–4, 2010. Proceedings, volume 6412 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16372-6 (print), 3-642-16373-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16373-9.

Peters:2010:TRSb

[PSS^{+}10b] James F. Peters, Andrzej Skowron, Roman Słowiński, Pawan Lingras, Duoqian Miao, and Shusaku Tsumoto, editors. Transactions on Rough Sets XII, volume 6190 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14466-7 (print), 3-642-14467-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14467-7.

Peters:2011:TRSb

[PSS^{+}11] James F. Peters, Andrzej Skowron, Hiroshi Sakai, Mihir Kumar Chakraborty, Dominik Slezak, Aboul Ella Hassanien, and William Zhu, editors. Transactions on Rough Sets XIV, volume 6600 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21562-9 (print), 3-642-21563-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21563-6.

Parent:2007:CMI

[PSST07] Christine Parent, Klaus-Dieter Schewe, Veda C. Storey, and Bernhard Thalheim, editors. Conceptual Modeling — ER 2007: 26th International Conference on Conceptual Modeling, Auckland, New Zealand, November 5–9, 2007. Proceedings, volume 4801 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75562-4 (print), 3-540-75563-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75563-0.

Paliouras:2011:KDM

[PST11] Georgios Paliouras, Constantine D. Spyropoulos, and George Tsatsaronis, editors. Knowledge-Driven Multimedia Information Extraction and Ontology Evolution: Bridging the Semantic Gap, volume 6050 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20794-4 (print), 3-642-20795-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20795-2.

Peters:2009:TRS

[PSW^{+}09] James F. Peters, Andrzej Skowron, Marcin Wolski, Mihir K. Chakraborty, and Wei-Zhi Wu, editors. Transactions on Rough Sets X, volume 5656 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03280-X (print), 3-642-03281-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03281-3.

Prehn:1991:VFSa

[PT91a] Søren Prehn and W. J.

Toetenel, editors. VDM ’91: Formal Software Development Methods: 4th International Symposium of VDM Europe, Noordwijkerhout, The Netherlands, October 21– 25, 1991: Proceedings. Part I. Conference Contributions, volume 551 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54834-3 (Berlin), 0-387-54834-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 V36 1991 v. 1–2 (c1991). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0551.
htm;   http://www.springerlink.
com/content/978-3-540-
54834-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=551.

Prehn:1991:VFSb

[PT91b] Søren Prehn and W. J. Toetenel, editors. VDM ’91: Formal Software Development Methods: 4th International Symposium of VDM Europe, Noordwijkerhout, The Netherlands, October 21–25, 1991: Proceedings. Part II. Tutorials, volume 552 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54868-8 (Berlin), 0-387-54868-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 V36 1991 v. 1–2 (c1991). URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0552.
htm;   http://www.springerlink.

com/content/978-3-540-
54868-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=552.

Pernul:1992:ERA

[PT92] G. Pernul and A. Min Tjoa, editors. Entityrelationship approach — ER ’92: 11th International Conference on the Entity-Relationship Approach, Karlsruhe, Germany, October 7– 9, 1992: proceedings, volume 645 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-56023-8 (Berlin), 0-387-56023-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26 I57 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0645.htm;
http://www.springerlink.
com/content/978-3-540-
56023-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=645.

Pacholski:1995:CSL

[PT95] Leszek Pacholski and Jerzy Tiuryn, editors. Computer science logic: 8th workshop, CSL ’94, Kazimierz, Poland, September 25–30, 1994: Selected Papers, volume 933 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60017-5 (Berlin), 0-387-60017-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W64 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0933.htm;
 http://www.springerlink.
com/content/978-3-540-
60017-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=933.

Peters:1997:RAT

[PT97] C. Peters and C. Thanos, editors. Research and advanced technology for digital libraries: First European Conference, ECDL ’97, Pisa, Italy, September 1–3, 1997: proceedings, volume 1324 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63554-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z692.C65 E28 1997. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t1324.htm;
http://www.springerlink.
com/content/978-3-540-
63554-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1324.

Pernici:1998:AIS

[PT98] Barbara Pernici and C. (Costantino) Thanos, editors. Advanced information systems engineering: 10th International Conference, CAiSE’98, Pisa, Italy, June 8– 12, 1998: Proceedings, volume 1413 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64556-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1413. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1413.htm;
http://www.springerlink.
com/content/978-3-540-
64556-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1413.

Pinter:1999:NGI

[PT99] Ron Yair Pinter and Shalom Tsur, editors. Next generation information technology and systems: 4th International Workshop, NGITS’99, Zikhron-Ya’akov, Israel, July 1999: Proceedings, volume 1649 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66225-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .N49 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1649.htm;
 http://www.springerlink.
com/content/978-3-540-
66225-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1649.

Peled:2005:ATV

[PT05] Doron A. Peled and Yih-Kuen Tsay, editors. Automated technology for verification and analysis: third international symposium, ATVA 2005, Taipei, Taiwan, October 4–7, 2005: proceedings, volume 3707 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29209-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I578 2005.

URL http://www.springerlink.

com/content/978-3-540-
29209-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3707.

Preneel:2006:SAC

[PT06a] Bart Preneel and Stafford Tavares, editors. Selected Areas in Cryptography: 12th International Workshop, SAC 2005, Kingston, ON, Canada, August 11–12, 2005, Revised Selected Papers, volume 3897 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33108-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33108-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3897.

Pugh:2006:LCP

[PT06b] William Pugh and Chau-Wen Tseng, editors. Languages and compilers for parallel computing: 15th workshop, LCPC 2002, College Park, MD, USA, July 25–27, 2002, revised papers, volume 2481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30781-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 2002. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        30781-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2481.

Pautasso:2008:SCI

[PT08a] Cesare Pautasso and Éric Tanter, editors. Software Composition: 7th International Symposium, SC 2008, Budapest, Hungary, March 29–30, 2008. Proceedings, volume 4954 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78788-7 (print), 3-540-78789-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78789-1.

Perron:2008:IAT

[PT08b] Laurent Perron and Michael A. Trick, editors. Integration of AI and OR Techniques in Constraint Programming for Combinatorial Optimization Problems: 5th International Conference, CPAIOR 2008 Paris, France, May 20–23, 2008 Proceedings, volume 5015 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68154-X (print), 3-540-68155-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68155-7.

Paletta:2009:ACS

[PT09] Lucas Paletta and John K.

Tsotsos, editors. Attention in Cognitive Systems: 5th International Workshop on Attention in Cognitive Systems, WAPCV 2008 Fira, Santorini, Greece, May 12, 2008 Revised Selected Papers, volume 5395 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00581-0 (print), 3-642-00582-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00582-4.

Preneel:2011:CHE

[PT11] Bart Preneel and Tsuyoshi Takagi, editors. Cryptographic Hardware and Embedded Systems — CHES 2011: 13th International Workshop, Nara, Japan, September 28 — October 1, 2011. Proceedings, volume 6917 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23950-1 (print), 3-642-23951-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23951-9.

Pavelka:1999:STP

[PTB99] Jan Pavelka, Gerard Tel, and Miroslav Bartosek, editors. SOF-SEM’99: theory and practice of informatics: 26th Conference on Current Trends in Theory and Practice of Informatics, Milovy, Czech Republic, November/ December, 1999: Proceedings, volume 1725 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66694-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1725; QA76.751 .S62 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1725.
htm;   http://www.springerlink.
com/content/978-3-540-
66694-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1725.

Pfoser:2011:AST

[PTM^{+}11] Dieter Pfoser, Yufei Tao, Kyriakos Mouratidis, Mario A. Nascimento, Mohamed Mokbel, Shashi Shekhar, and Yan Huang, editors. Advances in Spatial and Temporal Databases: 12th International Symposium, SSTD 2011, Minneapolis, MN, USA, August 24–26, 2011, Proceedings, volume 6849 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22921-2 (print), 3-642-22922-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22922-0.

Paletta:2005:APC

[PTRH05] Lucas Paletta, John K.

Tsotsos, Erich Rome, and Glyn Humphreys, editors. Attention and performance in computational vision: second international workshop, WAPCV 2004, Prague, Czech Republic, May 15, 2004. Revised selected papers, volume 3368 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24421-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .I69 2004. URL http:

//www.springerlink.
com/content/978-3-540-
24421-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3368;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105311.

Petta:2003:ESA

[PTZ03] Paolo Petta, Robert Tolksdorf, and Franco Zambonelli, editors. Engineering Societies in the Agents World III: Third International Workshop, ESAW 2002, Madrid, Spain, September 16– 17, 2002: Revised Papers, volume 2577 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-14009-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 E83 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2577.htm;
 http://www.springerlink.
com/content/978-3-540-
14009-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2577.

Petrenko:2004:FAS

[PU04] Alexandre Petrenko and Andreas Ulrich, editors. Formal Approaches to Software Testing: Third International Workshop on Formal Approaches to Testing of Software: FATES 2003: Montréal, Québec, Canada, October 6th, 2003: Revised Papers, volume 2931 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20894-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 F52 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2931.

                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20894-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2931;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95400.

Puebla:2007:LBP

[Pue07] Germán Puebla, editor. Logic-Based Program Synthesis and Transformation: 16th International Symposium, LOP-STR 2006, Venice, Italy, July 12–14, 2006, Revised Selected Papers, volume 4407 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71409-X (print), 3-540-71410-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71410-1.

Puppe:1999:XKB

[Pup99] Frank Puppe, editor. XPS-99: Knowledgebased systems: survey and future directions: 5th biannual German Conference on Knowledge-Based Systems, Würzburg, Germany, March 3–5, 1999: Proceedings, volume 1570 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65658-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 v.1570. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1570.htm;
 http://www.springerlink.
com/content/978-3-540-
65658-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1570.

Puppis:2010:ABL

[Pup10] Gabriele Puppis, editor.

Automata for Branching and Layered Temporal Structures: An Investigation into Regularities of Infinite Transition Systems, volume 5955 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CO-DEN LNCSD9. ISBN 3-642-11880-1 (print), 3-642-11881-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 170 (est.) pp. LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        11881-4.

Parigot:2000:LPA

[PV00] Michel Parigot and A. Voronkov, editors. Logic for Programming and Automated Reasoning: 7th International Conference, LPAR 2000, Reunion Island, France, November 6–10, 2000. Proceedings, volume 1955 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41285-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1955. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1955.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41285-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1955.

Peled:2002:FTN

[PV02] Doron A. Peled and Moshe Y. Vardi, editors. Formal Techniques for Networked and Distributed Systems— FORTE 2002: 22nd IFIP WG 6.1 International Conference, Houston, Texas, USA, November 11–14, 2002. Proceedings, volume 2529 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00141-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 F67 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2529.
htm;   http://www.springerlink.
com/content/978-3-540-
00141-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2529.

Pimentel:2004:CSA

[PV04] Andy D. Pimentel and Stamatis Vassiliadis, editors. Computer systems: architectures, modeling, and simulation: third and fourth international workshops, SAMOS 2003 and SAMOS 2004, Samos, Greece, July 21–23, 2003 and July 19–21, 2004: proceedings, volume 3133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22377-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A73 S28 2003.

URL http://www.springerlink.

                                        com/content/978-3-540-
                                        22377-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3133.

Paillier:2007:CHE

[PV07] Pascal Paillier and Ingrid Verbauwhede, editors. Cryptographic Hardware and Embedded Systems — CHES 2007: 9th International Workshop, Vienna, Austria, September 10–13, 2007. Proceedings, volume 4727 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74734-6 (print), 3-540-74735-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74735-2.

Poli:1999:EIA

[PVC^{+}99] Riccardo Poli, H.-M. Voigt, S. Cagnoni, D. Corne, G. Smith, and T. Fogarty, editors. Evolutionary image analysis, signal processing, and telecommunications: First European Workshops, EvoIASP ’99 and EuroEcTel ’99, Göteborg, Sweden, May 26–27, 1999: Proceedings, volume 1596 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65837-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5102.9 .E98 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1596.htm;
 http://www.springerlink.
com/content/978-3-540-
65837-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1596.

Pena:2012:FPA

[PvES12] Ricardo Peña, Marko van Eekelen, and Olha Shkaravska, editors. Foundational and Practical Aspects of Resource Analysis: Second International Workshop, FOPARA 2011, Madrid, Spain, May 19, 2011, Revised Selected Papers, volume 7177 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32494-0 (print), 3-642-32495-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        32495-6.

Petta:2005:MAS

[PVP05] Paolo Petta, László Zsolt Varga, and Michal Pěchouček, editors. Multi-agent systems and applications IV: 4th International Central and Eastern European Conference on Multi-agent Systems, CEEMAS 2005, Budapest, Hungary, September 15–17, 2005: proceedings, volume 3690 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29046-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 C43 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        29046-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3690.

Pras:2007:DAN

[PvS07] Aiko Pras and Marten van Sinderen, editors. Dependable and Adaptable Networks and Services: 13th Open European Summer School and IFIP TC6.6 Workshop, EUNICE 2007, Enschede, The Netherlands, July 18–20, 2007. Proceedings, volume 4606 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73529-1 (print), 3-540-73530-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73530-4.

Petrenko:2007:TSC

[PVTG07] Alexandre Petrenko, Margus Veanes, Jan Tretmans, and Wolfgang Grieskamp, editors. Testing of Software and Communicating Systems: 19th IFIP TC6/WG6.1 International Conference, TestCom 2007, 7th International Workshop, FATES 2007, Tallinn, Estonia, June 26–29, 2007. Proceedings, volume 4581 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73065-6 (print), 3-540-73066-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73066-8.

Paliouras:2005:ICS

[PVV05] Vassilis Paliouras, Johan Vounckx, and Diederik Verkest, editors. Integrated circuit and system design: power and timing modeling, optimization and simulation: 15th international workshop, PATMOS 2005, Leuven, Belgium, September 21–23, 2005: proceedings, volume 3728 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29013-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29013-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3728.

Pnueli:2010:PSI

[PVV10] Amir Pnueli, Irina Virbitskaite, and Andrei Voronkov, editors. Perspectives of Systems Informatics: 7th International Andrei Ershov Memorial Conference, PSI 2009, Novosibirsk, Russia, June 15–19, 2009. Revised Papers, volume 5947 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11485-7 (print), 3-642-11486-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11486-1.

Palanque:2010:HES

[PVW10] Philippe Palanque, Jean Vanderdonckt, and Marco Winckler, editors. Human Error, Safety and Systems Development: 7th IFIP WG 13.5 Working Conference, HESSD 2009, Brussels, Belgium, September 23–25, 2009, Revised Selected Papers, volume 5962 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11749-X (print), 3-642-11750-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11750-3.

Pearce:1992:LAE

[PW92a] David A. Pearce and G. Wagner, editors. Logics in AI: European Workshop JELIA ’92, Berlin, Germany, September 7–10, 1992: proceedings, volume 633 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55887-X (Berlin), 0-387-55887-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .E985 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0633.
htm;   http://www.springerlink.
com/content/978-3-540-
55887-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=633.

Pearce:1992:NLI

[PW92b] David A. Pearce and H. Wansing, editors. Nonclassical logics and information processing: International Workshop, Berlin, Germany, November 9–10, 1990: proceedings, volume 619 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55745-8 (Berlin), 0-387-55745-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.63.N658 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0619.htm;
 http://www.springerlink.
com/content/978-3-540-
55745-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=619.

“This volume comprises the proceedings of the First All-Berlin Workshop on Nonclassical Logics and Information Processing which was held at the Freie Universität on the first anniversary of the opening of the Berlin Wall, 9–10 November, 1990.”–Pref.

Petkov:2003:CAI

[PW03] Nicolai Petkov and Michel A.

Westenberg, editors. Computer Analysis of Images and Patterns: 10th International Conference, CAIP 2003, Groningen, The Netherlands, August 25–27, 2003: Proceedings, volume 2756 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40730-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .I535 2003. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2756.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40730-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2756;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13427.

Psaila:2007:CWT

[PW07] Giuseppe Psaila and Roland Wagner, editors. E-Commerce and Web Technologies: 8th International Conference, EC-Web 2007, Regensburg, Germany, September 3–7, 2007. Proceedings, volume 4655 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74562-9 (print), 3-540-74563-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74563-1.

Psaila:2008:CWT

[PW08] Giuseppe Psaila and Roland Wagner, editors. E-Commerce and Web Technologies: 9th International Conference, EC-Web 2008 Turin, Italy, September 3–4, 2008 Proceedings, volume 5183 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85716-8 (print), 3-540-85717-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85717-4.

Peled:2009:MCA

[PW09] Doron A. Peled and Michael J. Wooldridge, editors. Model Checking and Artificial Intelligence: 5th International Workshop, MoChArt 2008, Patras, Greece, July 21, 2008. Revised Selected and Invited Papers, volume 5348 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00430-X (print), 3-642-00431-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00431-5.

Preparata:2008:FAS

[PWY08] Franco P. Preparata, Xiaodong Wu, and Jianping Yin, editors. Frontiers in Algorithmics: Second Annual International Workshop, FAW 2008, Changsha, China, June 19–21, 2008, Proceedings, volume 5059 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69310-6 (print), 3-540-69311-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69311-6.

Pagnucco:1998:AMA

[PWZ98] Maurice Pagnucco, Wayne Wobcke, and Chengqi Zhang, editors. Agents and multi-agent systems: formalisms, methodologies and applications: based on the AI’97 Workshop on Commonsense Reasoning, Intelligent Agents, and Distributed Artificial Intelligence, Perth, Australia, December 1, 1997, volume 1441 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64769-4 (softcover).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A96 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1441.
htm;   http://www.springerlink.
com/content/978-3-540-
64769-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1441.

Patrick:2005:FCD

[PY05a] Andrew S. Patrick and Moti Yung, editors. Financial cryptography and data security: 9th international conference, FC 2005, Roseau, The Commonwealth of Dominica, February 28– March 3, 2005: revised papers, volume 3570 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26656-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 2005; HG1710 .F35 2005eb; HG1710; Internet. URL http:

//www.springerlink.
com/content/978-3-540-
26656-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3570.

Pettersson:2005:FMA

[PY05b] Paul Pettersson and Wang Yi, editors. Formal Modeling and Analysis of Timed Systems: Third International Conference, FOR-MATS 2005, Uppsala, Sweden, September 26– 28, 2005: Proceedings, volume 3829 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30946-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65 F676 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30946-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3829.

Pei:2008:ISC

[PYLW08] Dingyi Pei, Moti Yung, Dongdai Lin, and Chuankun Wu, editors. Information Security and Cryptology: Third SKLOIS Conference, Inscrypt 2007, Xining, China, August 31

— September 5, 2007, Revised Selected Papers, volume 4990 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79498-0 (print), 3-540-79499-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79499-8.

Papazoglou:1992:NGI

[PZ92] M. Papazoglou and J. Zeleznikow, editors. The Next generation of information systems: from data to knowledge, volume 611 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55616-8 (Berlin), 0-387-55616-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 N485 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0611.htm;
http://www.springerlink.
com/content/978-3-540-
55616-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=611.

“A selection of papers presented at two IJCAI-91 workshops, Sydney, Australia August 26, 1991.”.

Pedreschi:1996:LDI

[PZ96] Dino Pedreschi and Carlo Zaniolo, editors. Logic in databases: International Workshop LID ’96, San Miniato, Italy, July 1–2, 1996: proceedings, volume 1154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61814-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5857 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1154.htm;
 http://www.springerlink.
com/content/978-3-540-
61814-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1154.

Priami:2005:TCSb

[PZ05] Corrado Priami and Alexander Zelikovsky, editors. Transactions on Computational Systems Biology II, volume 3680 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29401-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH506 .T732 2005. URL http:

//www.springerlink.

com/content/978-3-540-
29401-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3680.

Padgham:2007:AOS

[PZ07a] Lin Padgham and Franco Zambonelli, editors. Agent-Oriented Software Engineering VII: 7th International Workshop, AOSE 2006, Hakodate, Japan, May 8, 2006, Revised and Invited Papers, volume 4405 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70944-4 (print), 3-540-70945-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70945-9.

Prencipe:2007:SIC

[PZ07b] Giuseppe Prencipe and Shmuel Zaks, editors. Structural Information and Communication Complexity: 14th International Colloquium, SIROCCO 2007, Castiglioncello, Italy, June 5–8, 2007. Proceedings, volume 4474 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72918-6 (print), 3-540-72951-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72951-8.

Papadimitriou:2008:INE

[PZ08] Christos Papadimitriou and Shuzhong Zhang, editors. Internet and Network Economics: 4th International Workshop, WINE 2008, Shanghai, China, December 17– 20, 2008. Proceedings, volume 5385 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92184-2 (print), 3-540-92185-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        92185-1.

Pieprzyk:2009:PST

[PZ09] Josef Pieprzyk and Fangguo Zhang, editors. Provable Security: Third International Conference, ProvSec 2009, Guangzhou, China, November 11–13, 2009. Proceedings, volume 5848 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04641-X (print), 3-642-04642-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04642-1.

Pan:2008:TLD

[PZE^{+}08] Zhigeng Pan, Xiaopeng Zhang, Abdennour El Rhalibi, Woontack Woo, and Yi Li, editors. Technologies for E-Learning and Digital Entertainment: Third International Conference, Edutainment 2008 Nanjing, China, June 25–27, 2008 Proceedings, volume 5093 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69734-9 (print), 3-540-69736-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69736-7.

Ponce:1996:ORC

[PZH96] Jean Ponce, Andrew Zisserman, and Martial Hebert, editors. Object representation in computer vision II: ECCV ’96 International Workshop, Cambridge, UK, April 13–14, 1996: proceedings, volume 1144 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61750-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .O255 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1144.htm;
 http://www.springerlink.
com/content/978-3-540-
61750-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1144.

Papadias:2007:AST

[PZK07] Dimitris Papadias, Donghui Zhang, and George Kollios, editors. Advances in Spatial and Temporal Databases: 10th International Symposium, SSTD 2007, Boston, MA, USA, July 16–18, 2007. Proceedings, volume 4605 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73539-9 (print), 3-540-73540-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73540-3.

Park:2012:NPC

[PZYS12] James J. Park, Albert Zomaya, Sang-Soo Yeo, and Sartaj Sahni, editors. Network and Parallel Computing: 9th IFIP International Conference, NPC 2012, Gwangju, Korea, September 6–8, 2012. Proceedings, volume 7513 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35605-2 (print), 3-642-35606-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35606-3.

Quaglini:2001:AIM

[QBA01] Silvana Quaglini, P. Barahona, and Steen Andreassen, editors. Artificial intelligence in medicine: 8th Conference on Artificial Intelligence in Medicine in Europe, AIME 2001, Cascais, Portugal, July 1–4, 2001: Proceedings, volume 2101 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42294-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R859.7.A78 C66 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2101.htm;
 http://www.springerlink.
com/content/978-3-540-
42294-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2101.

Quirchmayr:2012:MRP

[QBY^{+}12] Gerald Quirchmayr, Josef Basl, Ilsun You, Lida Xu, and Edgar Weippl, editors. Multidisciplinary Research and Practice for Information Systems: IFIP WG 8.4, 8.9/TC 5 International Cross-Domain Conference and Workshop on Availability, Reliability, and Security, CD-ARES 2012, Prague, Czech Republic, August 20–24, 2012. Proceedings, volume 7465 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32497-5 (print), 3-642-32498-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32498-7.

Quinonero-Candela:2006:MLC

[QCDM06] Joaquin Quinonero-Candela, Ido Dagan, and Bernardo Magnini, editors. Machine Learning Challenges: Evaluating Predictive Uncertainty, Visual Object Classification, and Recognising Tectual Entailment. First Pascal Machine Learning Challenges Workshop, MLCW 2005, Southampton, UK, April 11–13, 2005, Revised Selected Papers, volume 3944 of Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33427-0 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33427-9.

Qing:2003:ICS

[QGZ03] Sihan Qing, Dieter Gollmann, and Jianying Zhou, editors. Information and Communications Security: 5th International Conference: Proceedings/ICICS 2003, Huhehaote, China, October 10–13, 2003, volume 2836 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20150-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I554 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2836.
htm;   http://www.springerlink.
com/content/978-3-540-
20150-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2836;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13930.

Qin:2010:UTP

[Qin10] Shengchao Qin, editor. Unifying Theories of Programming: Third International Symposium, UTP 2010, Shanghai, China, November 15– 16, 2010. Proceedings, volume 6445 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16689-X (print), 3-642-16690-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16690-7.

Qing:2007:ICS

[QIW07] Sihan Qing, Hideki Imai, and Guilin Wang, editors. Information and Communications Security: 9th International Conference, ICICS 2007, Zhengzhou, China, December 12–15, 2007. Proceedings, volume 4861 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77047-X (print), 3-540-77048-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77048-0.

Qiu:2010:AMI

[QLK^{+}10] Guoping Qiu, Kin Man Lam, Hitoshi Kiya, Xiang-Yang Xue, C.- C. Jay Kuo, and Michael S. Lew, editors. Advances in Multimedia Information Processing — PCM 2010: 11th Pacific Rim Conference on Multimedia, Shanghai, China, September 21– 24, 2010, Proceedings, Part I, volume 6297 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15701-7 (print), 3-642-15702-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15702-8.

Qiu:2011:AMI

[QLK^{+}11] Guoping Qiu, Kin Man Lam, Hitoshi Kiya, Xiang-Yang Xue, C.- C. Jay Kuo, and Michael S. Lew, editors. Advances in Multimedia Information Processing — PCM 2010: 11th Pacific Rim Conference on Multimedia, Shanghai, China, September 2010, Proceedings, Part II, volume 6298 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-15695-9 (print), 3-642-15696-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15696-0.

Qiu:2007:AVI

[QLXL07] Guoping Qiu, Clement Leung, Xiangyang Xue, and Robert Laurini, editors. Advances in Visual Information Systems: 9th International Conference, VISUAL 2007 Shanghai, China, June 28–29, 2007 Revised Selected Papers, volume 4781 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76413-5 (print), 3-540-76414-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76414-4.

Qing:2005:ICS

[QMLW05] Sihan Qing, Wenbo Mao, Javier Lopez, and Guilin Wang, editors. Information and communications security: 7th international conference, ICICS 2005, Beijing, China, December 10–13, 2005: proceedings, volume 3783 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30934-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I554 2005.

URL http://www.springerlink.
com/content/978-3-540-
30934-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3783.

Qing:2009:ICS

[QMW09] Sihan Qing, Chris J. Mitchell, and Guilin Wang, editors. Information and Communications Security: 11th International Conference, ICICS 2009, Beijing, China, December 14– 17, 2009. Proceedings, volume 5927 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-11144-0 (print), 3-642-11145-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11145-7.

Qing:2001:ICS

[QOZ01] Sihan Qing, Tatsuaki Okamoto, and Jianying Zhou, editors. Information and Communications Security: Third International Conference, ICICS 2001, Xian, China, November 13– 16, 2001. Proceedings, volume 2229 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42880-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I554 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2229.htm;
http://www.springerlink.
com/content/978-3-540-
42880-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2229.

Qin:2011:FMS

[QQ11] Shengchao Qin and Zongyan Qiu, editors. Formal Methods and Software Engineering: 13th International Conference on Formal Engineering Methods, ICFEM 2011, Durham, UK, October 26–28, 2011. Proceedings, volume 6991 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24558-7 (print), 3-642-24559-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24559-6.

Quisquater:2000:SCR

[QS00] J.-J. Quisquater and Bruce Schneier, editors. Smart card research and applications: Third International Conference, CARDIS’98, Louvain-la-Neuve, Belgium, September 1998: Proceedings, volume 1820 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67923-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.S62 C36 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1820.
htm;   http://www.springerlink.
com/content/978-3-540-
67923-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1820.

Quirchmayr:1998:DES

[QSBC98] Gerald Quirchmayr, Erich Schweighofer, and T. J. M. Bench-Capon, editors. Database and Expert Systems Applications: 9th International Conference, DEXA ’98, Vienna, Austria, August 24–28, 1998: Proceedings, volume 1460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64950-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D15592 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1460.
htm;   http://www.springerlink.
com/content/978-3-540-
64950-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1460.

Qing:2011:ICS

[QSWL11] Sihan Qing, Willy Susilo, Guilin Wang, and Dongmei Liu, editors. Information and Communications Security: 13th International Conference, ICICS 2011, Beijing, China, November 23– 26, 2011. Proceedings, volume 7043 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25242-7 (print), 3-642-25243-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25243-3.

Quisquater:1998:CSE

[Qui98] J.-J. Quisquater, editor.

Computer security, ES-ORICS 98: 5th European Symposium on Research in Computer Security, Louvain-la-Neuve, Belgium, September 16– 18, 1998: Proceedings, volume 1485 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65004-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1485. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1485.htm;
 http://www.springerlink.
com/content/978-3-540-
65004-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1485.

Quintarelli:2004:MCB

[Qui04] Elisa Quintarelli. Model-Checking Based Data Retrieval: An Application to Semistructured and Temporal Data, volume 2917 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20971-9. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 134 pp. LCCN QA76.7 .Q49 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2917.htm;
 http://www.springerlink.
com/content/978-3-540-
20971-3;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2917;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94902.

Quisquater:1990:ACE

[QV90] Jean-Jacques Quisquater and Joos Vandewalle, editors. Advances in cryptology — EUROCRYPT ’89: Workshop on the Theory and Application of Cryptographic Techniques, Houthalen, Belgium, April 10–13, 1989: proceedings, volume 434 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-53433-4 (New York), 3-540-53433-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 1989; QA267.A1 L43 no.434. DM98.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0434.
htm;   http://www.springerlink.
com/content/978-0-387-
53433-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=434.

Reich:2000:OHS

[RA00] Siegfried Reich and Kenneth M. Anderson, editors. Open Hypermedia Systems and Structural Computing: 6th International Workshop, OHS-6, 2nd International Workshop, SC-2, San Antonio, Texas, USA, May 30–June 4, 2000. Proceedings, volume 1903 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41084-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I59 O37 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1903.htm;
 http://www.springerlink.
com/content/978-3-540-
41084-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1903.

Rashid:2006:TAOa

[RA06a] Awais Rashid and Mehmet Aksit. Transactions on aspect-oriented software development I, volume 3880 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32972-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

????  pp.  LCCN

???? URL http:

//www.springerlink.
com/content/978-3-540-
32972-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3880.

Rashid:2006:TAOb

[RA06b] Awais Rashid and Mehmet Aksit, editors. Transactions on Aspect-Oriented Software Development II, volume 4242 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48890-1 (print), 3-540-48891-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48891-0.

Rashid:2007:TAOa

[RA07a] Awais Rashid and Mehmet Aksit, editors. Transactions on Aspect-Oriented Software Development III, volume 4620 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-75161-0, 3-540-75162-9. vi + 199 pp. LCCN QA76.76.D47 T733 2007. URL

http://springerlink.
metapress.com/openurl.

                                        asp?genre=issue\%26issn=
                                        0302-9743\%26volume=
                                        4620;   http://www.
                                        springerlink.com/content/
                                        978-3-540-75161-8/.

Rashid:2007:TAOb

[RA07b] Awais Rashid and Mehmet Aksit, editors. Transactions on Aspect-Oriented Software Development IV, volume 4640 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77041-0 (print), 3-540-77042-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77042-8.

Ramsay:2012:AIM

[RA12] Allan Ramsay and Gennady Agre, editors. Artificial Intelligence: Methodology, Systems, and Applications: 15th International Conference, AIMSA 2012, Varna, Bulgaria, September 12–15, 2012. Proceedings, volume 7557 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33184-X (print), 3-642-33185-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33185-5.

Rabin:2010:ACC

[Rab10] Tal Rabin, editor. Advances in Cryptology — CRYPTO 2010: 30th Annual Cryptology Conference, Santa Barbara, CA, USA, August 15– 19, 2010. Proceedings, volume 6223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14622-8 (print), 3-642-14623-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14623-7.

Rist:2003:IVA

[RABR03] Thomas Rist, Ruth Aylett, Daniel Ballin, and Jeff Rickel, editors. Intelligent Virtual Agents: Fourth International Workshop, IVA 2003, Kloster Irsee, Germany, September 15– 17, 2003: Proceedings, volume 2792 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20003-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65 I93 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2792.
htm;   http://www.springerlink.
com/content/978-3-540-
20003-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2792;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12026.

Rahm:2004:DIL

[Rah04] Erhard Rahm, editor. Data Integration in the Life Sciences: First International Workshop, DILS 2004, Leipzig, Germany, March 25– 26, 2004: Proceedings, volume 2994 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21300-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .D54 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2994.htm;
 http://www.springerlink.
com/content/978-3-540-
21300-0;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2994;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96666.

Raidl:2003:AEC

[Rai03] Gunther Raidl, editor. Applications of evolutionary computing: EvoWorkshops 2003: Evo-BIO, EvoCOP, EvoIASP, EvoMUSART, EvoROB, and EvoSTIM, Essex, UK, April 14–16, 2003: Proceedings, volume 2611 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00976-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E899 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2611.htm;
http://www.springerlink.
com/content/978-3-540-
00976-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2611.

Rajsbaum:2002:LTI

[Raj02] Sergio Rajsbaum, editor. Latin 2002: theoretical informatics: 5th LATIN American symposium, Cancun, Mexico, April 3–6, 2002: Proceedings, volume 2286 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43400-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .L348 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2286.htm;
 http://www.springerlink.
com/content/978-3-540-
43400-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2286.

Rajasekaran:2009:BCB

[Raj09] Sanguthevar Rajasekaran, editor. Bioinformatics and Computational Biology: First International Conference, BI-CoB 2009, New Orleans, LA, USA, April 8–10, 2009. Proceedings, volume 5462 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00726-0 (print), 3-642-00727-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
00727-9.

Rall:1981:ADT

[Ral81]  Louis B. Rall.  Au-

tomatic differentiation: techniques and applications, volume 120 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CODEN LNCSD9. ISBN 0-387-10861-0, 3-540-10861-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 165 pp. LCCN QA304 .R34. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0120.htm;
 http://www.springerlink.
com/content/978-0-387-
10861-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=120.

Ralescu:1994:FLA

[Ral94] Anca L. Ralescu, editor. Fuzzy logic in artificial intelligence: IJCAI ’93 workshop, Chambery, France, August 28, 1993: proceedings, volume 847 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58409-9 (Berlin), 0-387-58409-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QC173.454 .F89 1994. DM39.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0847.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        58409-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=847.

Ramalingam:1996:BIC

[Ram96] G. Ramalingam. Bounded incremental computation, volume 1089 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61320-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 190 pp. LCCN QA76 .R263 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1089.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61320-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1089.

Raman:1998:AST

[Ram98] T. V. Raman. Audio system for technical readings, volume 1410 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65515-8. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 121 pp. LCCN QA267.A1 L43 no.1410. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1410.htm;
 http://www.springerlink.
com/content/978-3-540-
65515-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1410.

Originally presented as the author’s thesis (Ph. D.–Cornell University, 1994).

Ramakrishnan:2001:PAD

[Ram01] I. V. Ramakrishnan, editor. Practical aspects of declarative languages: Third International Symposium, PADL 2001, Las Vegas, Nevada, March 11–12, 2001: Proceedings, volume 1990 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41768-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .P32 2001; QA267.A1 L43 no.1990. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1990.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41768-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1990.

Richardson:2005:SPI

[RAM05] Ita Richardson, Pekka Abrahamsson, and Richard Messnarz, editors. Software process improvement: 12th European conference, EuroSPI 2005, Budapest, Hungary, November 9–11, 2005: proceedings, volume 3792 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30286-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 E772

                                        2005. URL http://www.
                                        loc.gov/catdir/toc/
                                        fy0604/2005935531.html;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        30286-5.

Ramalingam:2008:PLS

[Ram08] G. Ramalingam, editor. Programming Languages and Systems: 6th Asian Symposium, APLAS 2008, Bangalore, India, December 9– 11, 2008. Proceedings, volume 5356 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89329-6 (print), 3-540-89330-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89330-1.

Rangan:1993:NOS

[Ran93] P. Venkat Rangan, editor. Network and operating system support for digital audio and video: third International Workshop, La Jolla, California, USA, November 12–13, 1992: proceedings, volume 712 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57183-3, 0-387-57183-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7881.4 .N393 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0712.htm;
 http://www.springerlink.
com/content/978-3-540-
57183-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=712.

Raoult:1992:CCT

[Rao92] J.-C. Raoult, editor. CAAP ’92: 17th Colloquium on Trees in Algebra and Programming, Rennes, France, February 26–28, 1992, proceedings, volume 581 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55251-0 (Berlin), 0-387-55251-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166.2 .C65 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0581.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55251-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=581.

Rao:1995:EUM

[Rao95] Josyula Ramachandra Rao. Extensions of the UNITY methodology: compositionality, fairness, and probability in parallelism, volume 908 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 0-387-59173-7 (New York), 3-540-59173-7 (Berlin).

ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 178 pp. LCCN QA76.642 .R43 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0908.
htm;   http://www.springerlink.
com/content/978-0-387-
59173-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=908.

Rash:2001:FAA

[Ras01] James L. Rash, editor. Formal approaches to agent-based systems: First International Workshop, FAABS 2000, Greenbelt, MD, USA, April 5–7, 2000: Revised Papers, volume 1871 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42716-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1871. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1871.htm;
 http://www.springerlink.
com/content/978-3-540-
42716-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1871.

Rauchwerger:2004:LCP

[Rau04a] Lawrence Rauchwerger, editor. Languages and Compilers for Parallel Computing: 16th International Workshop, LCPC 2003, College Station, TX, October 2– 4, 2003, Revised Papers, volume 2958 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21199-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2958.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21199-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2958;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95707.

Rauterberg:2004:ECI

[Rau04b] Matthias Rauterberg, editor. Entertainment Computing—ICEC 2004: Third International Conference, Eindhoven, The Netherlands, September 1–3, 2004: Proceedings, volume 3166 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22947-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3166.htm;
 http://www.springerlink.
com/content/978-3-540-
22947-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3166;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99837.

Rau:2011:IDG

[Rau11] P. L. Patrick Rau, editor. Internationalization, Design and Global Development: 4th International Conference, IDGD 2011, Held as part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011. Proceedings, volume 6775 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21659-5 (print), 3-642-21660-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        21660-2.

Randall:2007:PAL

[RAW07] Marcus Randall, Hussein A. Abbass, and Janet Wiles, editors. Progress in Artificial Life: Third Australian Conference; ACAL 2007 Gold Coast, Australia, December 4–6, 2007 Proceedings, volume 4828 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76930-7 (print), 3-540-76931-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76931-6.

Richter:1991:PDK

[RB91] Michael M. Richter and H. Boley, editors. Processing declarative knowledge: international workshop, PDK ’91, Kaiserslautern, Germany, July 1–3, 1991: proceedings, volume 567 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-55033-X (Berlin), 0-387-55033-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .P7 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0567.htm;
 http://www.springerlink.
com/content/978-3-540-
55033-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=567.

Ruhe:2000:LSO

[RB00] Gunther Ruhe and Frank Bomarius, editors. Learning software organizations: methodology and applications: 11th International Conference on Software Engineering and Knowledge Engineering, SEKE’99, Kaiserslautern, Germany, June 16–19, 1999: Proceedings, volume 1756 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41430-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I573 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1756.htm;
 http://www.springerlink.
com/content/978-3-540-
41430-8;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1756.

Robshaw:2008:NSC

[RB08] Matthew Robshaw and Olivier Billet, editors. New Stream Cipher Designs: The eSTREAM Finalists, volume 4986 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68350-X (print), 3-540-68351-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68351-3.

Rome:2010:CII

[RB10] Erich Rome and Robin Bloomfield, editors. Critical Information Infrastructures Security: 4th International Workshop, CRITIS 2009, Bonn, Germany, September 30

— October 2, 2009. Revised Papers, volume 6027 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14378-4 (print), 3-642-14379-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14379-3.

Runarsson:2006:PPS

[RBB^{+}06] Thomas Philip Runarsson, Hans-Georg Beyer, Edmund Burke, Juan J. Merelo-Guervós, L. Darrell Whitley, and Xin Yao, editors. Parallel Problem Solving from Nature — PPSN IX: 9th International Conference, Reykjavik, Iceland, September 9–13, 2006, Proceedings, volume 4193 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38990-3 (print), 3-540-38991-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38991-0.

Rolland:1993:AIS

[RBC93] Colette Rolland, F. Bodart, and Corine Cauvet, editors. Advanced information systems engineering: 5th International Conference, CAiSE ’93, Paris, France, June 8– 11, 1993: proceedings, volume 685 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56777-1 (Berlin), 0-387-56777-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .N67 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0685.htm;
 http://www.springerlink.
com/content/978-3-540-
56777-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=685.

Rothlauf:2005:AEC

[RBC^{+}05] Franz Rothlauf, Jürgen Branke, Stefano Cagnoni, DavidW. Corne, Rolf Drechsler, Yaochu Jin, Penousal Machado, Elena Marchiori, Juan Romero, George D. Smith, and Giovanni Squillero, editors. Applications on evolutionary computing: EvoWorkshops 2005: EvoBIO, Evo-COMNET, EvoHOT, EvoIASP, EvoMUSART, and EvoSTOC, volume 3449 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25396-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.618 .E838 2005. URL http:

//www.springerlink.
com/content/978-3-540-
25396-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3449;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106856.

Rothlauf:2006:AEC

[RBC^{+}06] Franz Rothlauf, Jürgen Branke, Stefano Cagnoni, Ernesto Costa, Carlos Cotta, Rolf Drechsler, Evelyne Lutton, Penousal Machado, Jason H. Moore, Juan Romero, George D. Smith, Giovanni Squillero, and Hideyuki Takagi, editors. Applications of Evolutionary Computing: EvoWorkshops 2006: EvoBIO, Evo-COMNET, EvoHOT, EvoIASP, EvoINTER-ACTION, EvoMUSART, and EvoSTOC, Budapest, Hungary, April 10–12, 2006. Proceedings, volume 3907 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33237-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33237-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3907.

Renals:2006:MLM

[RBF06] Steve Renals, Samy Bengio, and Jonathan G. Fiscus, editors. Machine Learning for Multimodal Interaction: Third International Workshop, MLMI 2006, Bethesda, MD, USA, May 1–4, 2006, Revised Selected Papers, volume 4299 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69267-3 (print), 3-540-69268-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69268-3.

Reed:2009:SDM

[RBG09] Rick Reed, Attila Bilgic, and Reinhard Gotzhein, editors. SDL 2009: Design for Motes and Mobiles: 14th International SDL Forum Bochum, Germany, September 22– 24, 2009 Proceedings, volume 5719 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04553-7 (print), 3-642-04554-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04554-7.

Roth-Berghofer:2006:ACB

[RBGG06] Thomas R. Roth-Berghofer, Mehmet H. Göker, and H. Altay Güvenir, editors. Advances in Case-Based Reasoning: 8th European Conference, ECCBR 2006 Fethiye, Turkey, September 4– 7, 2006 Proceedings, volume 4106 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36843-4 (print), 3-540-36846-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36846-5.

Rasmussen:2004:PRD

[RBGS04] Carl Edward Rasmussen, Heinrich H. Bülthoff, Martin A. Giese, and Bernhard Schölkopf, editors. Pattern Recognition: 26th DAGM symposium, Tübingen, Germany, August 30– September 1, 2004: Proceedings, volume 3175 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22945-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3175.htm;
 http://www.springerlink.
com/content/978-3-540-
22945-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3175;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99676.

Rombach:1993:ESE

[RBS93] H. Dieter Rombach, Victor R. Basili, and Richard W. Selby, editors. Experimental software engineering issues: critical assessment and future directions: international workshop, Dagstuhl Castle, Germany, September 14–18, 1992: proceedings, volume 706 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57092-6 (Berlin), 0-387-57092-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .E993 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0706.htm;
 http://www.springerlink.
com/content/978-3-540-
57092-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=706.

Roddick:2006:ACM

[RBsC^{+}06] John F. Roddick, V. Richard Benjamins, Samira Si said Cherfi, Roger Chiang, Christophe Claramunt, Ramez A. Elmasri, Fabio Grandi, Hyoil Han, Martin Hepp, Miltiadis D. Lytras, Vojislav B. Mišić, Geert Poels, Il-Yeol Song, Juan Trujillo, and Christelle Vangenot, editors. Advances in Conceptual Modeling — Theory and Practice: ER 2006 Workshops BP-UML, CoMoGIS, COSS, ECDM, OIS, QoIS, SemWAT, Tucson, AZ, USA, November 6–9, 2006. Proceedings, volume 4231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47703-9 (print), 3-540-47704-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        47704-4.

Roth-Berghofer:2006:MRC

[RBSL06] Thomas R. Roth-Berghofer, Stefan Schulz, and David B. Leake, editors. Modeling and Retrieval of Context: Second International Workshop, MRC 2005, Edinburgh, UK, July 31–August 1, 2005, Revised Selected Papers, volume 3946 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33587-0 (softcover). LCCN QA765.758 .I577 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33587-0.

Rivera:2004:GCF

[RBTD04] Francisco Fernández Rivera, Marian Bubak, Andrés Gómez Tato, and Ramón Doallo, editors. Grid Computing: First European Across Grids Conference, Santiago de Compostela, Spain, February 13–14, 2003: Revised Papers, volume 2970 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21048-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 E97 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2970.
htm;   http://www.springerlink.
com/content/978-3-540-
21048-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2970;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95647.

Rollinger:1995:KAA

[RBW95] Claus-Rainer Rollinger, Wilfried Brauer, and Ipke Wachsmuth, editors. KI-95: advances in artificial intelligence: 19th Annual German Conference on Artificial Intelligence, Bielefeld, Germany, September 11–13, 1995: proceedings, volume 981 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60343-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G47 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0981.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60343-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=981.

Ramani:1990:KBC

[RCA90] S. Ramani, R. Chandrasekar, and K. S. R. Anjaneyulu, editors. Knowledge based computer systems: International Conference KBCS ’89, Bombay, India, December 11–13, 1989 proceedings, volume 444 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52850-4 (Berlin), 0-387-52850-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I567 1989. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0444.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        52850-0;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=444.

Raidl:2004:AEC

[RCB^{+}04] Günther R. Raidl, Stefano Cagnoni, Jürgen Branke, David W. Corne, Rolf Drechsler, Yaochu Jin, Colin G. Johnson, Penousal Machado, Elena Marchiori, Franz Rothlauf, George D. Smith, and Giovanni Squillero, editors. Applications of Evolutionary Computing: EvoWorkshops 2004: EvoBIO, EvoCOMNET, EvoHOT, EvoIASP, EvoMUSART, and EvoSTOC, Coimbra, Portugal, April 5– 7, 2004: proceedings, volume 3005 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21378-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E838 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3005.htm;
 http://www.springerlink.
com/content/978-3-540-
21378-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3005;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b96500.

Rahaman:2012:PVD

[RCC12] Hafizur Rahaman, Sanatan Chattopadhyay, and Santanu Chattopadhyay, editors. Progress in VLSI Design and Test: 16th International Symposium, VDAT 2012, Shibpur, India, July 1– 4, 2012. Proceedings, volume 7373 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31493-7 (print), 3-642-31494-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31494-0.

Rao:2008:DCN

[RCJ^{+}08] Shrisha Rao, Mainak Chatterjee, Prasad Jayanti, C. Siva Ram Murthy, and Sanjoy Kumar Saha, editors. Distributed Computing and Networking: 9th International Conference, ICDCN 2008, Kolkata, India, January 5–8, 2008. Proceedings, volume 4904 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77443-2 (print), 3-540-77444-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77444-0.

Rodriguez:2005:GSF

[RCLE05] M. Andrea Rodríguez, Isabel F. Cruz, Sergei Levashkin, and Max J. Egenhofer, editors. Geospatial Sematics: First International Conference, GEOS 2005, Mexico City, Mexico, November 29–30, 2005, Proceedings, volume 3799 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30288-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30288-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3799.

Romay:2010:HAI

[RCS10] Manuel Graña Romay, Emilio Corchado, and M. Teresa Garcia Sebastian, editors. Hybrid Artificial Intelligence Systems: 5th International Conference, HAIS 2010, San Sebastián, Spain, June 23– 25, 2010. Proceedings, Part I, volume 6076 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13768-7 (print), 3-642-13769-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13769-3.

Rauber:2005:RAT

[RCT05] Andreas Rauber, Stavros Christodoulakis, and A. Min Tjoa, editors. Research and advanced technology for digital libraries: 9th European conference, ECDL 2005, Vienna, Austria, September 18–23, 2005. proceedings, volume 3652 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28767-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28767-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3652.

Rembold:1984:MTC

[RD84] Ulrich Rembold and R. Dillmann, editors. Methods and tools for computer integrated manufacturing, volume 168 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-12926-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TS155.6 .A37 1983. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0168.htm;
http://www.springerlink.
com/content/978-0-387-
12926-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=168.

Rangan:2001:PCI

[RD01] C. Pandu Rangan and C. Ding, editors. Progress in Cryptology—INDOCRYPT 2001: Second International Conference on Cryptology in India, Chennai, India, December 16–20, 2001. Proceedings, volume 2247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-43010-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I5535 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2247.
htm;   http://www.springerlink.
com/content/978-3-540-
43010-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2247.

Regnell:2012:REF

[RD12a] Björn Regnell and Daniela Damian, editors. Requirements Engineering: Foundation for Software Quality: 18th International Working Conference, REFSQ 2012, Essen, Germany, March 19–22, 2012. Proceedings, volume 7195 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28713-1 (print), 3-642-28714-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28714-5.

Roychoudhury:2012:TAC

[RD12b] Abhik Roychoudhury and Meenakshi D’Souza, editors. Theoretical Aspects of Computing — ICTAC 2012: 9th International Colloquium, Bangalore, India, September 24–27, 2012. Proceedings, volume 7521 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32942-X (print), 3-642-32943-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32943-2.

Rudolph:2009:CSL

[RDK09] Sebastian Rudolph, Frithjof Dau, and Sergei O. Kuznetsov, editors. Conceptual Structures: Leveraging Semantic Technologies: 17th International Conference on Conceptual Structures, ICCS 2009, Moscow, Russia, July 26–31, 2009. Proceedings, volume 5662 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03078-5 (print), 3-642-03079-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        03079-6.

Romanovsky:2001:AEH

[RDKT01] Alexander Romanovsky, Christophe Dony, Jørgen Lindskov Knudsen, and Anand Tripathi, editors. Advances in Exception Handling Techniques, volume 2022 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41952-7. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 288 pp. LCCN QA76.76.D47 A375 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2022.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41952-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2022.

Real:2011:CAIa

[RDPMA^{+}11a] Pedro Real, Daniel Diaz-Pernil, Helena Molina-Abril, Ainhoa Berciano, and Walter Kropatsch, editors. Computer Analysis of Images and Patterns: 14th International Conference, CAIP 2011, Seville, Spain, August 29–31, 2011, Proceedings, Part I, volume 6854 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23671-5 (print), 3-642-23672-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23672-3.

Real:2011:CAIb

[RDPMA^{+}11b] Pedro Real, Daniel Diaz-Pernil, Helena Molina-Abril, Ainhoa Berciano, and Walter Kropatsch, editors. Computer Analysis of Images and Patterns: 14th International Conference, CAIP 2011, Seville, Spain, August 29–31, 2011, Proceedings, Part II, volume 6855 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23677-4 (print), 3-642-23678-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23678-5.

Read:2001:ADB

[Rea01] Brian Read, editor. Advances in databases: 18th British National Conference on Databases, BN-COD 18, Chilton, UK, July 19–11, 2001: Proceedings, volume 2097 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42265-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B75 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2097.htm;
 http://www.springerlink.
com/content/978-3-540-
42265-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2097.

Reulke:2006:CIA

[REF^{+}06] Ralf Reulke, Ulrich Eckardt, Boris Flach, Uwe Knauer, and Konrad Polthier, editors. Combinatorial Image Analysis: 11th International Workshop, IWCIA 2006, Berlin, Germany, June 19–21, 2006. Proceedings, volume 4040 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35153-1 (print), 3-540-35154-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35154-2.

Reif:1988:VAA

[Rei88] J. H. Reif, editor. VLSI algorithms and architectures: 3rd Aegean Workshop on Computing, AWOC 88, Corfu, Greece, June 28–July 1, 1988: proceedings, volume 319 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-96818-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874 .A3381 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0319.htm;
http://www.springerlink.
com/content/978-0-387-
96818-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=319.

Reinfrank:1989:NMR

[Rei89] M. Reinfrank, editor.

Non-monotonic reasoning: 2nd International Workshop, Grassau, FRG, June 1988: proceedings, volume 346 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 3-540-50701-9 (Berlin), 0-387-50701-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .N66 1989.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0346.
htm;   http://www.springerlink.
com/content/978-3-540-
50701-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=346.

Papers presented at Second International Workshop on Non-Monotonic Reasoning, held in June 1988 in Grassau, West Germany.

Reiss:1993:RPO

[Rei93] Thomas H. Reiss. Recognizing planar objects using invariant image features, volume 676 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56713-5 (New York), 3-540-56713-5 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 180 pp. LCCN TA1650 .R45 1993. DM46.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0676.

htm;   http://www.springerlink.
com/content/978-0-387-
56713-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=676.

Reinelt:1994:TSC

[Rei94] G. (Gerhard) Reinelt.

The traveling salesman: computational solutions for TSP applications, volume 840 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58334-3 (Berlin), 0-387-58334-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 223 pp. LCCN T57.78 .R47 1994. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0840.
htm;   http://www.springerlink.
com/content/978-3-540-
58334-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=840.

Reichel:1995:FCT

[Rei95] Horst Reichel, editor.

Fundamentals of computation theory: 10th international conference, FCT ’95, Dresden, Germany, August 1995: proceedings, volume 965 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60249-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .F367 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0965.htm;
 http://www.springerlink.
com/content/978-3-540-
60249-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=965.

Reinartz:1999:FSD

[Rei99] Thomas Reinartz. Focusing Solutions for Data Mining: Analytical Studies and Experimental Results in Real-World Domains, volume 1623 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66429-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 307 pp. LCCN QA76.9.D343 R45 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1623.
htm;   http://www.springerlink.
com/content/978-3-540-
66429-1;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1623.

Reijers:2003:DCW

[Rei03] Hajo A. Reijers. Design and Control of Workflow Processes: Business Process Management for the Service Industry, volume 2617 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-01186-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 320 pp. LCCN HD62.17 .R45 2003. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2617.
htm;   http://www.springerlink.
com/content/978-3-540-
01186-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2617.

Reingold:2009:TCT

[Rei09] Omer Reingold, editor.

Theory of Cryptography: 6th Theory of Cryptography Conference, TCC 2009, San Francisco, CA, USA, March 15– 17, 2009. Proceedings, volume 5444 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00456-3 (print), 3-642-00457-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00457-5.

Renz:2002:QSR

[Ren02] Jochen Renz. Qualitative Spatial Reasoning with Topological Information, volume 2293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43346-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 207 pp. LCCN Q387 .R5 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2293.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43346-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2293.

Rosca:2006:ICA

[REPH06] Justinian Rosca, Deniz Erdogmus, José C. Príncipe, and Simon Haykin, editors. Independent Component Analysis and Blind Signal Separation: 6th International Conference, ICA 2006, Charleston, SC, USA, March 5–8, 2006. Proceedings, volume 3889 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32630-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32630-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3889.

Retore:1997:LAC

[Ret97] Christian Retoré, editor. Logical aspects of computational linguistics: First International Conference, LACL ’96, Nancy, France, September 23–25, 1996: Selected Papers, volume 1328 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63700-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN P98.5.L63L33 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1328.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        63700-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1328.

Reusch:1997:CIT

[Reu97] B. Reusch, editor. Computational intelligence: theory and applications: international conference, 5th fuzzy days, Dortmund, Germany, April 28–30 1997: proceedings, volume 1226 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62868-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q342 .C67 1997.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1226.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        62868-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1226.

Reusch:1999:CIT

[Reu99] B. Reusch, editor. Computational intelligence: theory and applications: International Conference, 6th fuzzy days, Dortmund, Germany, May 25–28, 1999: Proceedings, volume 1625 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66050-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q342 .C67 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1625.
htm;   http://www.springerlink.
com/content/978-3-540-
66050-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1625.

Reusch:2001:CIT

[Reu01] B. Reusch, editor. Computational intelligence: theory and applications: International Conference, 7th Fuzzy Days, Dortmund, Germany, October 1–3, 2001: Proceedings, volume 2206 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42732-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2206; Q342 .D67 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2206.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        42732-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2206.

Reyes:2004:CFM

[Rey04] Carlos Andrés Peña Reyes. Coevolutionary Fuzzy Modeling, volume 3204 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22994-9. ISSN 0302-9743 (print), 1611-3349 (electronic). 121 pp. LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3204.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22994-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3204;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99869.

Rizzo:1999:NGC

[RF99a] Luigi Rizzo and Serge Fdida, editors. Networked group communication: First International COST264 Workshop, NGC’99, Pisa, Italy, November 17– 20, 1999: Proceedings, volume 1736 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66782-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1736.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1736.
htm;   http://www.springerlink.
com/content/978-3-540-
66782-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1736.

Rudolph:1999:JSS

[RF99b] Larry Rudolph and Dror G. Feitelson, editors. Job scheduling strategies for parallel processing: IPPS/SPDP’99 workshop, JSSPP’99, San Juan, Puerto Rico, April 16, 1999: Proceedings, volume 1659 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66676-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 J757 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1659.htm;
 http://www.springerlink.
com/content/978-3-540-
66676-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1659.

Radig:2001:PRD

[RF01] Bernd Radig and Stefan Florczyk, editors. Pattern recognition: 23rd DAGM Symposium, Munich, Germany, September 12–14, 2001: Proceedings, volume 2191 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42596-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1650 .D35 2001; QA267.A1 L43 no.2191. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2191.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42596-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2191.

Rahman:2010:WAC

[RF10] Md. Saidur Rahman and Satoshi Fujita, editors. WALCOM: Algorithms and Computation: 4th International Workshop, WALCOM 2010, Dhaka, Bangladesh, February 10–12, 2010. Proceedings, volume 5942 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11439-3 (print), 3-642-11440-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11440-3.

Rosner:2012:CNL

[RF12] Michael Rosner and Norbert E. Fuchs, editors. Controlled Natural Language: Second International Workshop, CNL 2010, Marettimo Island, Italy, September 13–15, 2010. Revised Papers, volume 7175 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31174-1 (print), 3-642-31175-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31175-8.

Rothermel:2009:QCF

[RFBD09] Kurt Rothermel, Dieter Fritsch, Wolfgang Blochinger, and Frank Dürr, editors. Quality of Context: First International Workshop, QuaCon 2009, Stuttgart, Germany, June 25–26, 2009. Revised Papers., volume 5786 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04558-8 (print), 3-642-04559-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04559-2.

Ralyte:2012:AIS

[RFBW12] Jolita Ralyté, Xavier Franch, Sjaak Brinkkemper, and Stanislaw Wrycza, editors. Advanced Information Systems Engineering: 24th International Conference, CAiSE 2012, Gdansk, Poland, June 25–29, 2012. Proceedings, volume 7328 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31094-X (print), 3-642-31095-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31095-9.

Rangarajan:2003:EMM

[RFZ03] Anand Rangarajan, Mário Figueiredo, and Josiane Zerubia, editors. Energy Minimization Methods in Computer Vision and Pattern Recognition: 4th International Workshop, EMMCVPR 2003, Lisbon, Portugal, July 7–9, 2003: Proceedings, volume 2683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40498-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E54 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2683.htm;
http://www.springerlink.
com/content/978-3-540-
40498-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2683.

Renals:2003:TST

[RG03] Steve Renals and Gregory Grefenstette, editors. Text- and Speech-Triggered Information Access: 8th ELSNET Summer School Chios Island, Greece, July 15– 30, 2000 Revised Lectures, volume 2705 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40635-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.I58 E97 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2705.
htm;   http://www.springerlink.
com/content/978-3-540-
40635-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2705;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11786.

Raidl:2005:ECC

[RG05] Günther R. Raidl and Jens Gottlieb, editors. Evolutionary computation in combinatorial optimization: 5th European conference, EVO-COP 2005, Lausanne, Switzerland, March 30– April 1, 2005: Proceedings, volume 3448 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25337-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 E96 2005. URL http:

//www.springerlink.

com/content/978-3-540-
25337-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3448;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107115.

Rudolph:2011:WRR

[RG11] Sebastian Rudolph and Claudio Gutierrez, editors. Web Reasoning and Rule Systems: 5th International Conference, RR 2011, Galway, Ireland, August 29–30, 2011. Proceedings, volume 6902 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23579-4 (print), 3-642-23580-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23580-1.

Ruiz:2009:AHM

[RGLA09] Pedro M. Ruiz and Jose Joaquin Garcia-Luna-Aceves, editors. Ad-Hoc, Mobile and Wireless Networks: 8th International Conference, ADHOC-NOW 2009, Murcia, Spain, September 22–25, 2009 Proceedings, volume 5793 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04382-8 (print), 3-642-04383-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04383-3.

Richter-Gebert:2001:ADG

[RGW01] Jurgen Richter-Gebert and Dongming Wang, editors. Automated deduction in geometry: Third International Workshop, ADG 2000, Zürich, Switzerland, September 25–27, 2000: Revised Papers, volume 2061 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42598-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA448.D38 I577 2000; QA267.A1 L43 no.2061. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2061.htm;
 http://www.springerlink.
com/content/978-3-540-
42598-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2061.

Rothermel:1998:MAS

[RH98] Kurt Rothermel and Fritz Hohl, editors. Mobile agents: Second International Workshop, MA ’98, Stuttgart, Germany, September 9– 11, 1998: Proceedings, volume 1477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64959-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58M3 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1477.htm;
http://www.springerlink.
com/content/978-3-540-
64959-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1477.

Roddick:2001:TSS

[RH01] John F. Roddick and Kathleen Hornsby, editors. Temporal, spatial, and spatio-temporal data mining: First International Workshop, TSDM 2000, Lyon, France, September 12, 2000: Revised Papers, volume 2007 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41773-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 T73 2000; QA267.A1 L43 no.2007. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2007.htm;
 http://www.springerlink.
com/content/978-3-540-
41773-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2007.

Rome:2008:TAB

[RHD08] Erich Rome, Joachim Hertzberg, and Georg Dorffner, editors. Towards Affordance-Based Robot Control: International Seminar, Dagstuhl Castle, Germany, June 5–9, 2006. Revised Papers, volume 4760 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77914-0 (print), 3-540-77915-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
77915-5.

Riano:2008:KMH

[Ria08] David Riaño, editor.

Knowledge Management for Health Care Procedures: From Knowledge to Global Care, AIME 2007 Workshop K4CARE 2007, Amsterdam, The Netherlands, July 7, 2007, Revised Selected Papers, volume 4924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78623-6 (print), 3-540-78624-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78624-5.

Riano:2009:KMH

[Ria09] David Riaño, editor.

Knowledge Management for Health Care Procedures: ECAI 2008 Workshop, K4HelP 2008, Patras, Greece, July 21, 2008, Revised Selected Papers, volume 5626 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03261-3 (print), 3-642-03262-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03262-2.

Richter:1998:ALT

[Ric98] Michael M. Richter, editor. Algorithmic learning theory: 9th International Conference, ALT ’98, Otzenhausen, Germany, October 8–10, 1998: Proceedings, volume 1501 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65013-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 Lr3 no.1501. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1501.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65013-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1501.

Richardson:2002:MTR

[Ric02] Stephen D. Richardson, editor. Machine translation: from research to real users: 5th Conference of the Association for Machine Translation in the Americas, AMTA 2002, Tiburon, CA, USA, October 8– 12, 2002: Proceedings, volume 2499 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44282-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN P308 .A85 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2499.htm;
 http://www.springerlink.
com/content/978-3-540-
44282-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2499.

Rigoll:2008:PRD

[Rig08] Gerhard Rigoll, editor. Pattern Recognition: 30th DAGM Symposium Munich, Germany, June 10–13, 2008 Proceedings, volume 5096 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69320-3 (print), 3-540-69321-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69321-5.

Rahman:2012:WAC

[RiN12] Md. Saidur Rahman and Shin ichi Nakano, editors. WALCOM: Algorithms and Computation: 6th International Workshop, WAL-COM 2012, Dhaka, Bangladesh, February 15–17, 2012. Proceedings, volume 7157 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28075-7 (print), 3-642-28076-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28076-4.

Rudolph:2008:PPS

[RJL^{+}08] Günter Rudolph, Thomas Jansen, Simon Lucas, Carlo Poloni, and Nicola Beume, editors. Parallel Problem Solving from Nature — PPSN X: 10th International Conference Dortmund, Germany, September 13–17, 2008 Proceedings, volume 5199 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87699-5 (print), 3-540-87700-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87700-4.

Rong:2008:ATC

[RJS^{+}08] Chunming Rong, Martin Gilje Jaatun, Frode Eika Sandnes, Laurence T. Yang, and Jianhua Ma, editors. Autonomic and Trusted Computing: 5th International Conference, ATC 2008, Oslo, Norway, June 23– 25, 2008 Proceedings, volume 5060 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69294-0 (print), 3-540-69295-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69295-9.

Ras:1993:MIS

[RK93] Zbigniew Ras and J. Komorowski, editors. Methodologies for intelligent systems: 7th International Symposium, IS-MIS ’93, Trondheim, Norway, June 15–18, 1993: proceedings, volume 689 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56804-2 (Berlin), 0-387-56804-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0689.htm;
 http://www.springerlink.
com/content/978-3-540-
56804-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=689.

Roli:2002:MCS

[RK02] Fabio Roli and Josef Kittler, editors. Multiple Classifier Systems: Third International Workshop, MCS 2002, Cagliari, Italy, June 24–26, 2002. Proceedings, volume 2364 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43818-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .M84 2002. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t2364.htm;
http://www.springerlink.
com/content/978-3-540-
43818-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2364.

Proceedings of the Third International Workshop on Multiple Classifier Systems (MCS 2002), held at Grand Hotel Chia Laguna, Cagliari, Italy, June 24–26, 2002.

Reimer:2006:PAK

[RK06] Ulrich Reimer and Dimitris Karagiannis, editors. Practical Aspects of Knowledge Management: 6th International Conference, PAKM 2006, Vienna, Austria, November 30 — December1, 2006. Proceedings, volume 4333 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49998-9 (print), 3-540-49999-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49999-2.

Richards:2009:KAA

[RK09] Debbie Richards and Byeong-Ho Kang, editors. Knowledge Acquisition: Approaches, Algorithms and Applications: Pacific Rim Knowledge Acquisition Workshop, PKAW 2008, Hanoi, Vietnam, December 15–16, 2008, Revised Selected Papers, volume 5465 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01714-2 (print), 3-642-01715-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01715-5.

Richards:2012:KMA

[RK12] Deborah Richards and Byeong Ho Kang, editors. Knowledge Management and Acquisition for Intelligent Systems: 12th Pacific Rim Knowledge Acquisition Workshop, PKAW 2012, Kuching, Malaysia, September 5– 6, 2012. Proceedings, volume 7457 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32540-8 (print), 3-642-32541-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32541-0.

Romer:2006:WSN

[RKM06] Kay Römer, Holger Karl, and Friedemann Mattern, editors. Wireless sensor networks: third European workshop, EWSN 2006, Zürich, Switzerland, February 13–15, 2006: proceedings, volume 3868 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32158-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7872.D48 E94 2006.

URL http://www.springerlink.
com/content/978-3-540-
32158-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3868.

Rautiainen:2012:GPC

[RKM^{+}12] Mika Rautiainen, Timo Korhonen, Edward Mutafungwa, Eila Ovaska, Artem Katasonov, Antti Evesti, Heikki Ailisto, Aaron Quigley, Jonna Häkkilä, Natasa Milic-Frayling, and Jukka Riekki, editors. Grid and Pervasive Computing Workshops: International Workshops, S3E, HWTS, Doctoral Colloquium, Held in Conjunction with GPC 2011, Oulu, Finland, May 11–13, 2011. Revised Selected Papers, volume 7096 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27915-5 (print), 3-642-27916-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27916-4.

Ruttkay:2009:IVA

[RKNV09] Zsófia Ruttkay, Michael Kipp, Anton Nijholt, and Hannes Högni Vilhjálmsson, editors. Intelligent Virtual Agents: 9th International Conference, IVA 2009 Amsterdam, The Netherlands, September 14–16, 2009 Proceedings, volume 5773 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04379-8 (print), 3-642-04380-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
04380-2.

Rafanelli:1989:SSD

[RKS89] M. Rafanelli, J. C.

Klensin, and P. Svensson, editors. Statistical and scientific database management: fourth International Working Conference SSDBM, Rome, Italy, June 21–23, 1988, 15, volume 339 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 3-540-50575-X, 0-387-50575-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.339. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0339.htm;
http://www.springerlink.
com/content/978-3-540-
50575-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=339.

Rutkowski:2012:AISa

[RKS^{+}12a] Leszek Rutkowski, Marcin Korytkowski, Rafał Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M. Zurada, editors. Artificial Intelligence and Soft Computing: 11th International Conference, ICAISC 2012, Zakopane, Poland, April 29–May 3, 2012, Proceedings, Part I, volume 7267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29346-8 (print), 3-642-29347-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29347-4.

Rutkowski:2012:AISb

[RKS^{+}12b] Leszek Rutkowski, Marcin Korytkowski, Rafał Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M. Zurada, editors. Artificial Intelligence and Soft Computing: 11th International Conference, ICAISC 2012, Zakopane, Poland, April 29–May 3, 2012, Proceedings, Part II, volume 7268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29349-2 (print), 3-642-29350-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29350-4.

Rutkowski:2012:SEC

[RKS^{+}12c] Leszek Rutkowski, Marcin Korytkowski, Rafał Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M. Zurada, editors. Swarm and Evolutionary Computation: International Symposia, SIDE 2012 and EC 2012, Held in Conjunction with ICAISC 2012, Zakopane, Poland, April 29–May 3, 2012. Proceedings, volume 7269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29352-2 (print), 3-642-29353-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29353-5.

Roli:2004:MCS

[RKW04] Fabio Roli, Josef Kittler, and Terry Windeatt, editors. Multiple Classifier Systems: 5th International Workshop, MCS 2004, Cagliari, Italy, June 9–11, 2004: Proceedings, volume 3077 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22144-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5.I574 2004. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3077.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        22144-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3077;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98227.

Rhee:2006:ISC

[RL06] Min Surp Rhee and Byoungcheon Lee, editors. Information Security and Cryptology — ICISC 2006: 9th International Conference, Busan, Korea, November 30 — December 1, 2006. Proceedings, volume 4296 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49112-0 (print), 3-540-49114-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49114-9.

Rizzo:2011:CIM

[RL11a] Riccardo Rizzo and Paulo J. G. Lisboa, editors. Computational Intelligence Methods for Bioinformatics and Biostatistics: 7th International Meeting, CIBB 2010, Palermo, Italy, September 16–18, 2010, Revised Selected Papers, volume 6685 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21945-4 (print), 3-642-21946-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21946-7.

Rocha:2011:PAD

[RL11b] Ricardo Rocha and John Launchbury, editors. Practical Aspects of Declarative Languages: 13th International Symposium, PADL 2011, Austin, TX, USA, January 24–25, 2011. Proceedings, volume 6539 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18377-8 (print), 3-642-18378-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18378-2.

Ravenscroft:2012:CLC

[RLKHL12]  Andrew  Ravenscroft,

Stefanie Lindstaedt, Carlos Delgado Kloos, and Davinia Hernández-Leo, editors. 21st Century Learning for 21st Century Skills: 7th European Conference of Technology Enhanced Learning, EC-TEL 2012, Saarbrücken, Germany, September 18– 21, 2012. Proceedings, volume 7563 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33262-5 (print), 3-642-33263-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33263-0.

Ramamoorthy:2004:SER

[RLL04] C. V. Ramamoorthy, Roger Lee, and Kyung Whan Lee, editors. Software Engineering Research and Applications: First International Conference, SERA 2003, San Francisco, CA, USA, June 25–27, 2003: Selected Revised Papers, volume 3026 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21975-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E965 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3026.
htm;   http://www.springerlink.
com/content/978-3-540-
21975-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3026;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97161.

Roggen:2008:SSC

[RLT^{+}08] Daniel Roggen, Clemens Lombriser, Gerhard Tröster, Gerd Kortuem, and Paul Havinga, editors. Smart Sensing and Context: Third European Conference, EuroSSC 2008, Zurich, Switzerland, October 29–31, 2008. Proceedings, volume 5279 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88792-X (print), 3-540-88793-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88793-5.

Ras:1996:FIS

[RM96] Zbigniew Ras and Maciej Michalewicz, editors. Foundations of intelligent systems: 9th international symposium, ISMIS ’96, Zakopane, Poland, June 9–13, 1996: proceedings, volume 1079 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61286-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1079.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61286-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1079.

Papers selected for presentation at the Ninth International Symposium on Methodologies for Intelligent Systems.

Ralescu:1997:FLA

[RM97a] Anca L. Ralescu and T. P. Martin, editors.

Fuzzy logic in artificial intelligence: towards intelligent systems: IJCAI ’95 workshop, Montréal, Canada, August 19– 21, 1995, Selected Papers, volume 1188 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62474-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q335 .F93 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1188.htm;
 http://www.springerlink.
com/content/978-3-540-
62474-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1188.

Reischuk:1997:SAS

[RM97b] Rudiger Reischuk and Michel Morvan, editors. STACS 97: 14th Annual Symposium on Theoretical Aspects of Computer Science, Lübeck, Germany, February 27– March 1, 1997: proceedings, volume 1200 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62616-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S956 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1200.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62616-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1200.

Ranchhod:2002:ANL

[RM02] Elisabete Ranchhod and Nuno J. Mamede, editors. Advances in Natural Language Processing: Third International Conference, PorTAL 2002, Faro, Portugal, June 23–26, 2002. Proceedings, volume 2389 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43829-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 P67 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2389.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43829-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2389.

PorTAL—Portugal for Natural Language Processing— follows two previous conferences: FracTAL, held at the Université de Franche-Comte, Besançon (France), in December 1997, and Vex-TAL, held at Venice International University, Ca Foscari (Italy), in November 1999.

Ribeiro:2004:EEA

[RM04a] Celso C. Ribeiro and Simone L. Martins, editors. Experimental and Efficient Algorithms: Third International Workshop, WEA 2004, Angra Dos Reis, Brazil, May 25– 28, 2004, Proceedings, volume 3059 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22067-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3059.htm;
 http://www.springerlink.
com/content/978-3-540-
22067-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3059;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b97914.

Roy:2004:FSE

[RM04b] Bimal Roy and Willi Meier, editors. Fast Software Encryption: 11th International Workshop, FSE 2004: Delhi, India, February 5–7, 2004: Revised Papers, volume 3017 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22171-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3017.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22171-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3017;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98177.

Rahwan:2009:AMA

[RM09] Iyad Rahwan and Pavlos Moraitis, editors. Argumentation in Multi-Agent Systems: Fifth International Workshop, ArgMAS 2008, Estoril, Portugal, May 12, 2008. Revised Selected and Invited Papers, volume 5384 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00206-4 (print), 3-642-00207-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00207-6.

Ricciato:2010:TMA

[RMB10] Fabio Ricciato, Marco Mellia, and Ernst Biersack, editors. Traffic Monitoring and Analysis: Second International Workshop, TMA 2010, Zurich, Switzerland, April 7, 2010.Proceedings, volume 6003 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12364-3 (print), 3-642-12365-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12365-8.

Rajaraman:2010:DCS

[RMDS10] Rajmohan Rajaraman, Thomas Moscibroda, Adam Dunkels, and Anna Scaglione, editors. Distributed Computing in Sensor Systems: 6th IEEE International Conference, DCOSS 2010, Santa Barbara, CA, USA, June 21–23, 2010. Proceedings, volume 6131 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13650-8 (print), 3-642-13651-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13651-1.

Ryan:2004:OAF

[RME04] D. Ryan, John-Jules Ch. Meyer, and Hans-Dieter Ehrich, editors. Objects, Agents, and Features: International Seminar, Dagstuhl Castle, Germany, February 16–21, 2003, Revised and Invited Papers, volume 2975 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21989-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
21989-7;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2975;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97772.

Raubal:2006:GIS

[RMFG06] Martin Raubal, Harvey J. Miller, Andrew U. Frank, and Michael F. Goodchild, editors. Geographic Information Science: 4th International Conference, GIScience 2006, Münster, Germany, September 20– 23, 2006. Proceedings, volume 4197 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-44526-9 (print), 3-540-44528-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
44528-9.

Rueda:2007:PPR

[RMK07] Luis Rueda, Domingo Mery, and Josef Kittler, editors. Progress in Pattern Recognition, Image Analysis and Applications: 12th Iberoamericann Congress on Pattern Recognition, CIARP 2007, Valparaiso, Chile, November 13–16, 2007. Proceedings, volume 4756 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76724-X (print), 3-540-76725-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76725-1.

Rahwan:2005:AMA

[RMR05] Iyad Rahwan, Pavlos Moraitis, and Chris Reed, editors. Argumentation in multiagent systems: first international workshop, ArgMAS 2004, New York, NY, USA, July 19, 2004. Revised selected and invited papers, volume 3366 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24526-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 A73 2004. URL http:

//www.springerlink.
com/content/978-3-540-
24526-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3366;

http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105815.

Rattray:2004:AMS

[RMS04] C. (Charles) Rattray, Savitri Maharaj, and Carron Shankland, editors. Algebraic methodology and software technology: 10th international conference, AMAST 2004, Stirling, Scotland, UK, July 12–16, 2004: proceedings, volume 3116 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22381-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I5713 2004. URL http:

//www.springerlink.
com/content/978-3-540-
22381-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3116;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98770.

Reussner:2005:QSA

[RMS^{+}05] Ralf Reussner, Johannes Mayer, Judith A. Stafford, Sven Overhage, Steffen Becker, and Patrick J. Schroeder, editors. Quality of software architectures and software quality: First International Conference on the Quality of Software Architectures, QoSA 2005 and Second International Workshop on Software Quality, SOQUA 2005, Erfurt, Germany, September 20–22, 2005: proceedings, volume 3712 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29033-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.754 .Q35 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29033-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3712.

Rofer:2012:RRS

[RMSS12] Thomas Röfer, N. Michael Mayer, Jesus Savage, and Uluc? Saranlı, editors. RoboCup 2011: Robot Soccer World Cup XV, volume 7416 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32059-7 (print), 3-642-32060-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32060-6.

Rinderle-Ma:2011:BPM

[RMTW11] Stefanie Rinderle-Ma, Farouk Toumani, and Karsten Wolf, editors. Business Process Management: 9th International Conference, BPM 2011, Clermont-Ferrand, France, August 30 — September 2, 2011. Proceedings, volume 6896 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23058-X (print), 3-642-23059-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23059-2.

Roberts:1996:BNT

[RMV96] James Roberts, Ugo Mocci, and Jorma Virtamo. Broadband network teletraffic: performance evaluation and design of broadband multiservice networks: final report of action COST 242, volume 1155 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61815-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xxii + 584 pp. LCCN TK5103.4 .B765 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1155.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61815-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1155.

Rhee:2011:ISC

[RN11] Kyung-Hyune Rhee and DaeHun Nyang, editors. Information Security and Cryptology — ICISC 2010: 13th International Conference, Seoul, Korea, December 1–3, 2010, Revised Selected Papers, volume 6829 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24208-1 (print), 3-642-24209-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24209-0.

Ras:2000:FIS

[RO00a] Zbigniew W. Raś and Setsuo Ohsuga, editors. Foundations of Intelligent Systems: 12th International Symposium, ISMIS 2000, Charlotte, NC, USA, October 11– 14, 2000. Proceedings, volume 1932 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41094-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1932. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1932.htm;
 http://www.springerlink.
com/content/978-3-540-
41094-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1932.

Roy:2000:PCI

[RO00b] Bimal Roy and Eiji Okamoto, editors. Progress in Cryptology—INDOCRYPT 2000: First International Conference in Cryptology in India, Calcutta, India, December 10–13, 2000. Proceedings, volume 1977 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41452-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1977.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1977.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        41452-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1977.

Rashid:2009:TAO

[RO09] Awais Rashid and Harold Ossher, editors. Transactions on Aspect-Oriented Software Development V, volume 5490 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02058-5 (print), 3-642-02059-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02059-9.

Robinet:1974:PSP

[Rob74] B. Robinet, editor. Programming Symposium: proceedings. Colloque sur la programmation, Paris, April 9–11, 1974, volume 19 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 0-387-06859-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75171 1974. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0019.
htm;   http://www.springerlink.
com/content/978-0-387-
06859-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=19.

French or English.

Robinet:1980:ISP

[Rob80] B. Robinet, editor. International Symposium on Programming: proceedings of the fourth Colloque international sur la programmation, Paris, 22–24 April, 1980, volume 83 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-09981-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I575 1980. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0083.htm;
 http://www.springerlink.
com/content/978-0-387-
09981-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=83.

English or French.

Roberto:1993:IPS

[Rob93] V. Roberto. Intelligent perceptual systems: new directions in computational perception, volume 745 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-57379-8. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 378 pp. LCCN QA76.76.E95 I5653 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0745.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57379-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=745.

Robshaw:2006:FSE

[Rob06] Matthew Robshaw, editor. Fast Software Encryption: 13th International Workshop, FSE 2006, Graz, Austria, March 15–17, 2006, Revised Selected Papers, volume 4047 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36597-4 (print), 3-540-36598-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36598-3.

Robertson:2011:EHA

[Rob11] Michelle M. Robertson, editor. Ergonomics and Health Aspects of Work with Computers: International Conference, EHAWC 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011. Proceedings, volume 6779 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21715-X (print), 3-642-21716-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21716-6.

Rocha:1992:NNT

[Roc92] A. F. (Armando Freitas) Rocha. Neural nets: a theory for brains and machines, volume 638 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55949-3 (Berlin), 0-387-55949-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 393 pp. LCCN QA76.87 .R63 1992. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0638.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        55949-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=638.

Rocca:2007:TLC

[Roc07] Simona Ronchi Della Rocca, editor. Typed Lambda Calculi and Applications: 8th International Conference,TLCA 2007, Paris, France,June 26–28, 2007. Proceedings, volume 4583 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73227-6 (print), 3-540-73228-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73228-0.

Rogaway:2011:ACC

[Rog11] Phillip Rogaway, editor. Advances in Cryptology — CRYPTO 2011: 31st Annual Cryptology Conference, Santa Barbara, CA, USA, August 14–18, 2011. Proceedings, volume 6841 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22791-0 (print), 3-642-22792-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22792-9.

Rolim:1997:RAT

[Rol97] Jose D. P. Rolim, editor.

Randomization and approximation techniques in computer science: International Workshop RANDOM’97, Bologna, Italy, July 11–12, 1997: proceedings, volume 1269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63248-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76..R266 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1269.
htm;   http://www.springerlink.
com/content/978-3-540-

                                        63248-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1269.

Rolim:1998:PDP

[Rol98] Jose D. P. Rolim, editor. Parallel and distributed processing: 10 IPPS/SPDP ’98 workshops held in conjunction with the 12th International Parallel Processing Symposium and 9th Symposium on Parallel and Distributed Processing, Orlando, Florida, USA, March 30–April 3, 1998: Proceedings, volume 1388 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64359-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P368 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1388.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64359-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1388.

Rolim:1999:PDP

[Rol99] Jose D. P. Rolim, editor. Parallel and distributed processing: 11th IPPS/ SPDP ’99 workshops held in conjunction with the 13th International Parallel Processing Symposium and 10th Symposium on Parallel and Distributed Processing, San Juan, Puerto Rico, USA, April 12–16, 1999: Proceedings, volume 1586 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65831-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P368 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1586.htm;
 http://www.springerlink.
com/content/978-3-540-
65831-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1586.

Rolim:2000:PDP

[Rol00] Jose D. P. Rolim, editor. Parallel and distributed processing: 15 IPDPS 2000 workshops, Cancun, Mexico, May 1–5, 2000: Proceedings, volume 1800 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67442-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58. P3686 2000; QA267.A1 L43 no.1800. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1800.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67442-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1800.

Ronse:1984:FSR

[Ron84] Christian Ronse. Feedback shift registers, volume 169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-13330-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). 145 pp. LCCN TK7895.S54 R66 1984. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0169.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        13330-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=169.

Rossi:2003:PPC

[Ros03] Francesca Rossi, editor. Principles and Practice of Constraint Programming: CP 2003:

9th International Conference, CP 2003, Kinsale, Ireland, September 29– October 3, 2003: Proceedings, volume 2833 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20202-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C7 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2833.htm;
http://www.springerlink.
com/content/978-3-540-
20202-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2833;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13743.

Rovan:1990:MFC

[Rov90] B. Rovan, editor. Mathematical foundations of computer science, 1990: Banská Bystrica, Czechoslovakia, August 27–31, 1990: proceedings, volume 452 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52953-5, 0-387-52953-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 M386 1990.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0452.
htm;   http://www.springerlink.
com/content/978-3-540-
52953-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=452.

Papers selected for presentation at the 15th Symposium on Mathematical Foundations of Computer Science — MFCS ’90.

Rovan:1998:STP

[Rov98] Branislav Rovan, editor. SOFSEM ’98: Theory and Practice of Informatics: 25th Conference on Current Trends in Theory and Practice of Informatics, Jasná, Slovakia, November 21–27, 1998. Proceedings, volume 1521 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65260-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .S62 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1521.htm;
 http://www.springerlink.
com/content/978-3-540-
65260-1;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1521.

Royer:1987:CTP

[Roy87]  James S. Royer.  A

connotational theory of program structure, volume 273 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-18253-5 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). 186 pp. LCCN QA76.7 .R69 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0273.htm;
http://www.springerlink.
com/content/978-0-387-
18253-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=273.

A revision of author’s thesis (Ph. D. — State University of New York at Buffalo).

Roy:2005:ACA

[Roy05] Bimal Roy, editor. Advances in cryptology: ASIACRYPT 2005: 11th International Conference on the Theory and Application of Cryptology and Information Security, Chennai, India, December 4–8, 2005, volume 3788 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30684-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I555 2005.

URL http://www.springerlink.
com/content/978-3-540-
30684-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3788.

Rozenberg:1986:APN

[Roz86] Grzegorz Rozenberg, editor. Advances in Petri nets, 1985, volume 222 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16480-4 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.222. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0222.htm;
 http://www.springerlink.
com/content/978-0-387-
16480-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=222.

“Covers the 6th European Workshop on Applications and Theory of Petri Nets held in Espoo, Finland in June 1985”– Pref.

Rozenberg:1987:APN

[Roz87] Grzegorz Rozenberg, editor. Advances in Petri nets, 1987, volume 266 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18086-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.266. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0266.htm;
 http://www.springerlink.
com/content/978-0-387-
18086-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=266.

Rozenberg:1988:APN

[Roz88] Grzegorz Rozenberg, editor. Advances in Petri nets, 1988, volume 340 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-50580-6 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.340. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0340.htm;

                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        50580-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=340.

Rozenberg:1990:APN

[Roz90] Grzegorz Rozenberg, editor. Advances in Petri nets, 1989, volume 424 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52494-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.424. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0424.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        52494-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=424.

Rozenberg:1991:APNa

[Roz91] Grzegorz Rozenberg, editor. Advances in Petri nets, 1990, volume 483 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-53863-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A48 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0483.htm;
http://www.springerlink.
com/content/978-0-387-
53863-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=483.

“Covers the 10th ‘International Conference on Applications and Theory of Petri Nets’, held in Bonn, Germany in June 1989”–Pref.

Rozenberg:1992:APN

[Roz92] Grzegorz Rozenberg, editor. Advances in Petri nets, 1992, volume 609 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55610-9, 0-387-55610-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.609. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0609.htm;
http://www.springerlink.
com/content/978-3-540-
55610-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=609.

Rozenberg:1993:APN

[Roz93] Grzegorz Rozenberg, editor. Advances in Petri Nets, 1993, volume 674 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-56689-9, 3-540-56689-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.674. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0674.htm;
 http://www.springerlink.
com/content/978-0-387-
56689-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=674.

Robert:2006:HPC

[RPBP06] Yves Robert, Manish Parashar, Ramamurthy Badrinath, and Viktor K. Prasanna, editors. High Performance Computing

— HiPC 2006: 13th International Conference, Bangalore, India, December 18–21, 2006. Proceedings, volume 4297 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68039-X (print), 3-540-68040-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68040-6.

Rahwan:2008:AMA

[RPR08] Iyad Rahwan, Simon Parsons, and Chris Reed, editors. Argumentation in Multi-Agent Systems: 4th International Workshop, ArgMAS 2007, Honolulu, HI, USA, May 15, 2007, Revised Selected and Invited Papers, volume 4946 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78914-6 (print), 3-540-78915-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78915-4.

Rothermel:1997:MAF

[RPZ97] Kurt Rothermel and R. Popescu-Zeletin, editors. Mobile agents: First International Workshop, MA ’97, Berlin, Germany, April 7–8, 1997: proceedings, volume 1219 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62803-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M3 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1219.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62803-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1219.

Robbins:1989:CXM

[RR89] Kay A. Robbins and Steven Robbins. The Cray X-MP/Model 24: a case study in pipelined architecture and vector processing, volume 374 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-97089-4. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 165 pp. LCCN QA76.8.C72 R631 1989. US$18.90. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0374.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        97089-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=374.

Rusinowitch:1993:CTR

[RR93] M. Rusinowitch and J. L. Remy, editors. Conditional term rewriting systems, volume 656 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56393-8 (Berlin), 0-387-56393-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I578 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0656.htm;
 http://www.springerlink.
com/content/978-3-540-
56393-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=656.

Ravn:1998:FTR

[RR98a] Anders P. Ravn and Hans Rischel, editors. Formal techniques in real-time and fault-tolerant systems: 5th International Symposium, FTRTFT ’98, Lyngby, Denmark, September 14–18, 1998: Proceedings, volume 1486 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65003-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54.F75 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1486.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        65003-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1486.

Reisig:1998:LPNa

[RR98b] Wolfgang Reisig and Grzegorz Rozenberg, editors. Lectures on Petri nets: advances in Petri nets. Part 1. Basic Models, volume 1491 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65306-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .L39 1998 1–2 (c1998). URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1491.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65306-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1491.

“Based on the Advanced Course on Petri Nets, held in Dagstuhl (Germany) in September 1996”.

Reisig:1998:LPNb

[RR98c] Wolfgang Reisig and Grzegorz Rozenberg, editors. Lectures on Petri nets: advances in Petri nets. Part 2. Applications, volume 1492 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65307-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .L39 1998 1–2 (c1998). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1492.htm;
 http://www.springerlink.
com/content/978-3-540-
65307-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1492.

“Based on the Advanced Course on Petri Nets, held in Dagstuhl (Germany) in September 1996”.

Reed:2001:SMU

[RR01] Rick Reed and Jeanne Reed, editors. SDL 2001: meeting UML, 10th International SDL Forum Copenhagen, Denmark, June 27–29, 2001: Proceedings, volume 2078 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42281-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.S35 S35 2001; QA267.A1 L43 no.2078. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2078.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42281-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2078.

Reed:2003:SSD

[RR03] Rick Reed and Jeanne Reed, editors. SDL 2003: system design: 11th international SDL Forum, Stuttgart, Germany, July 1–4, 2003: Proceedings, volume 2708 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40539-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.S35 S35 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2708.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40539-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2708.

Regin:2004:IAT

[RR04a] Jean-Charles Régin and Michel Rueher, editors. Integration of AI and OR Techniques in Constraint Programming for Combinatorial Optimization Problems, volume 3011 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21836-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3011.htm;
http://www.springerlink.
com/content/978-3-540-
21836-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3011;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96957.

Roca:2004:IMN

[RR04b] Vincent Roca and Franck Rousseau, editors. Interactive Multimedia and Next Generation Networks: Second International Workshop on Multimedia Interactive Protocols and Systems, MIPS 2004, Grenoble, France, November 16– 19, 2004. Proceedings, volume 3311 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23928-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.15 .I6 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3311.
htm;   http://www.springerlink.
com/content/978-3-540-
23928-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3311;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b103343.

Ramakrishnan:2008:TAC

[RR08] C. R. Ramakrishnan and Jakob Rehof, editors. Tools and Algorithms for the Construction and Analysis of Systems: 14th International Conference, TACAS 2008, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2008, Budapest, Hungary, March 29–April 6, 2008. Proceedings, volume 4963 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78799-2 (print), 3-540-78800-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78800-3.

Ramanujam:2012:DCI

[RR12] R. Ramanujam and Srini Ramaswamy, editors. Distributed Computing and Internet Technology: 8th International Conference, ICDCIT 2012, Bhubaneswar, India, February 2–4, 2012. Proceedings, volume 7154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28072-2 (print), 3-642-28073-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28073-3.

Rauch:2009:FIS

[RRBE09] Jan Rauch, Zbigniew W.

Raś, Petr Berka, and Tapio Elomaa, editors. Foundations of Intelligent Systems: 18th International Symposium, ISMIS 2009, Prague, Czech Republic, September 14–17, 2009. Proceedings, volume 5722 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04124-8 (print), 3-642-04125-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04125-9.

Rozenberg:1985:APN

[RRG85] Grzegorz Rozenberg, G. Roucairol, and Hartmann Jochen Genrich, editors. Advances in Petri nets, 1984, volume 188 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CODEN LNCSD9. ISBN 3-540-15204-0, 0-387-15204-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.188.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0188.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        15204-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=188.

Papers from two sessions of the European Workshop on Applications and Theory of Petri Nets, the 4th held in Toulouse in 1983 and the 5th held in Aarhus in 1984.

Richardson:2006:SPI

[RRM06] Ita Richardson, Per Runeson, and Richard Messnarz, editors. Software Process Improvement: 13th European Conference, EuroSPI 2006, Joensuu, Finland, October 11–13, 2006. Proceedings, volume 4257 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47695-4 (print), 3-540-47696-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47696-2.

Rausch:2008:CCM

[RRMP08] Andreas Rausch, Ralf Reussner, Raffaela Mirandola, and František Plášil, editors. The Common Component Modeling Example: Comparing Software Component Models, volume 5153 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85288-3 (print), 3-540-85289-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85289-6.

Rozenberg:1991:APNb

[RRP91] Grzegorz Rozenberg, Wolfgang Reisig, and H. Plunnecke, editors. Advances in Petri nets 1991, volume 524 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54398-8 (Berlin), 0-387-54398-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.524. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0524.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54398-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=524.

“Covers the 11th ‘International Conference on Applications and Theory of Petri Nets’ held in Paris, France in June 1990”–P. [v] Includes “Bibliography of Petri Nets 1990” by H. Plunnecke and W. Reisig: p. [317]-572.

Raman:1999:FST

[RRR99] V. Raman, C. Pandu Rangan, and R. Ramanujam, editors. Foundations of Software Technology and Theoretical Computer Science: 19th Conference, Chennai, India, December 13–15, 1999: Proceedings, volume 1738 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66836-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1738. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1738.htm;
 http://www.springerlink.
com/content/978-3-540-
66836-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1738.

Restivo:2001:TCS

[RRR01] Antonio Restivo, Luca Roversi, and S. Ronchi Della Rocca, editors. Theoretical computer science: 7th Italian Conference, ICTCS 2001, Torino, Italy, October 4–6, 2001: Proceedings, volume 2202 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42672-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 I28 2001. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2202.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42672-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2202.

Ramos:2005:ADS

[RRU05] Felix F. Ramos, Victor Larios Rosillo, and Herwig Unger, editors. Advanced Distributed Systems: 5th International School and Symposium, ISSADS 2005, Guadalajara, Mexico, January 24–28, 2005, Revised Selected Papers, volume 3563 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28063-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28063-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3563.

Rozenberg:1974:S

[RS74] Grzegorz Rozenberg and Arto Salomaa, editors. L systems, volume 15 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 0-387-06867-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH491 .L2 1974. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0015.htm;
 http://www.springerlink.
com/content/978-0-387-
06867-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=15.

Most of the papers were presented at a conference in Aarhus, Denmark, Jan. 14–25, 1974.

Riloff:1996:CSS

[RS96] Ellen Riloff and Gabriele Scheler. Connectionist, statistical, and symbolic approaches to learning for natural language processing, volume 1040 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60925-3. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 468 pp. LCCN QA76.9.N38 C67 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1040.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        60925-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1040.

Ramesh:1997:FST

[RS97a] S. Ramesh and G. Sivakumar, editors. Foundations of software technology and theoretical computer science: 17th conference, Kharagpur, India, December 18– 20, 1997: proceedings, volume 1346 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63876-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1346.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1346.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63876-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1346.

Ras:1997:FIS

[RS97b] Zbigniew Ras and Andrzej Skowron, editors. Foundations of intelligent systems: 10th international symposium, ISMIS ’97, Charlotte, North Carolina, USA, October 15–18, 1997: proceedings, volume 1325 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63614-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1325.
htm;   http://www.springerlink.
com/content/978-3-540-
63614-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1325.

Ralescu:1999:FLA

[RS99a] Anca L. Ralescu and James G. Shanahan, editors. Fuzzy logic in artificial intelligence: IJ-CAI’97 Workshop, Nagoya, Japan, August 23–24, 1997: Selected and Invited Papers, volume 1566 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66374-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .F93 1999.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t1566.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66374-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1566.

Ras:1999:FIS

[RS99b] Zbigniew Ras and Andrzej Skowron, editors. Foundations of intelligent systems: 11th International Symposium, ISMIS’99, Warsaw, Poland, June 8– 11, 1999: Proceedings, volume 1609 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65965-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1609.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65965-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1609.

Rao:2000:NGN

[RS00] S. Rao and Kaare Ingar Sletta, editors. Next Generation Networks.

Networks and Services for the Information Society: 5th IFIP TC6 International Symposium, INTERWORKING 2000, Bergen, Norway, October 3–6, 2000. Proceedings, volume 1938 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41140-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1938. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1938.htm;
 http://www.springerlink.
com/content/978-3-540-
41140-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1938.

Rouveirol:2001:ILP

[RS01] Celine Rouveirol and Michele Sebag, editors. Inductive logic programming: 11th International Conference, ILP 2001, Strasbourg, France, September 9–11, 2001: Proceedings, volume 2157 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42538-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I52 2001; QA267.A1 L43 no.2157. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2157.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42538-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2157.

Rosen:2003:RST

[RS03] Jean-Pierre Rosen and Alfred Strohmeier, editors. Reliable software technologies — Ada-Europe 2003: 8th Ada-Europe International Conference on Reliable Software Technologies, Toulouse, France, June 16–20, 2003: Proceedings, volume 2655 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40376-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2655.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40376-0;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2655.

Rao:2005:CHE

[RS05] Josyula R. Rao and Berk Sunar, editors. Cryptographic Hardware and Embedded Systems — CHES 2005: 7th International Workshop, Edinburgh, UK, August 29– September 1, 2005. Proceedings, volume 3659 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28474-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28474-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3659.

Ramanujam:2006:FFS

[RS06] R. Ramanujam and Sandeep Sen, editors. FSTTCS 2005: Foundations of Software Technology and Theoretical Computer Science: 25th International Conference, Hyderabad, India, December 15–18, 2005. Proceedings, volume 3821 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30495-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        30495-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3821.

Ramanujam:2009:LAT

[RS09a] R. Ramanujam and Sundar Sarukkai, editors. Logic and Its Applications: Third Indian Conference, ICLA 2009, Chennai, India, January 7–11, 2009. Proceedings, volume 5378 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-92700-X (print), 3-540-92701-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        92701-3.

Roedig:2009:WSN

[RS09b] Utz Roedig and Cormac J. Sreenan, editors. Wireless Sensor Networks: 6th European Conference, EWSN 2009, Cork, Ireland, February 11–13, 2009. Proceedings, volume 5432 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00223-4 (print), 3-642-00224-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00224-3.

Roy:2009:PCI

[RS09c] Bimal Roy and Nicolas Sendrier, editors. Progress in Cryptology — INDOCRYPT 2009: 10th International Conference on Cryptology in India, New Delhi, India, December 13–16, 2009. Proceedings, volume 5922 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10627-7 (print), 3-642-10628-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10628-6.

Ryan:2009:VIS

[RS09d] Peter Y. A. Ryan and Berry Schoenmakers, editors. E-Voting and Identity: Second International Conference, VOTE-ID 2009, Luxembourg, September 7– 8, 2009. Proceedings, volume 5767 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04134-5 (print), 3-642-04135-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04135-8.

Raman:2010:PEC

[RS10] Venkatesh Raman and Saket Saurabh, editors. Parameterized and Exact Computation: 5th International Symposium, IPEC 2010, Chennai, India, December 13– 15, 2010. Proceedings, volume 6478 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17492-2 (print), 3-642-17493-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17493-3.

Richa:2012:SSS

[RS12] Andréa W. Richa and Christian Scheideler, editors. Stabilization, Safety, and Security of Distributed Systems: 14th International Symposium, SSS 2012, Toronto, Canada, October 1–4, 2012. Proceedings, volume 7596 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33535-7 (print), 3-642-33536-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33536-5.

Reps:2007:PAC

[RSB07] Thomas Reps, Mooly Sagiv, and Jörg Bauer, editors. Program Analysis and Compilation, Theory and Practice: Essays Dedicated to Reinhard Wilhelm on the Occasion of His 60th Birthday, volume 4444 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71315-8 (print), 3-540-71322-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        71322-7.

Ryan:2003:GPE

[RSK03] Conor Ryan, Terence Soule, and Maarten Keijzer, editors. Genetic programming: 6th European Conference, EuroGP 2003, Essex, UK, April 14–16, 2003: Proceedings, volume 2610 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00971-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.623 .E9 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2610.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00971-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2610.

Ruiz-Shulcloper:2008:PPR

[RSK08] José Ruiz-Shulcloper and Walter G. Kropatsch, editors. Progress in Pattern Recognition, Image Analysis and Applications: 13th Iberoamerican Congress on Pattern Recognition, CIARP 2008, Havana, Cuba, September 9–12, 2008. Proceedings, volume 5197 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85919-5 (print), 3-540-85920-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85920-8.

Robertson:2001:SAS

[RSL01] Paul Robertson, Howard E.

Shrobe, and Robert Laddaga, editors. Selfadaptive software: First International Workshop, IWSAS 2000, Oxford, UK, April 17–19, 2000: Revised Papers, volume 1936 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41655-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.S375 I97 2000; QA267.A1 L43 no.1936. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1936.htm;
 http://www.springerlink.
com/content/978-3-540-
41655-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1936.

Reussner:2006:AST

[RSS06] Ralf H. Reussner, Judith A. Stafford, and Clemens A. Szyperski, editors. Architecting Systems with Trustworthy Components: International Seminar, Dagstuhl Castle, Germany, December 12–17, 2004. Revised Selected Papers, volume 3938 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35800-5 (print), 3-540-35833-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        35833-6.

Reichl:2009:NEN

[RST09] Peter Reichl, Burkhard Stiller, and Bruno Tuffin, editors. Network Economics for Next Generation Networks: 6th International Workshop on Internet Charging and Qos Technologies, ICQT 2009, Aachen, Germany, May 11–15, 2009. Proceedings, volume 5539 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01795-9 (print), 3-642-01796-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01796-4.

Rutkowski:2010:AISb

[RST^{+}10a] Leszek Rutkowski, Rafał Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M. Zurada, editors. Artifical Intelligence and Soft Computing: 10th International Conference, ICAISC 2010, Zakopane, Poland, June 13–17, 2010, Part II, volume 6114 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13231-6 (print), 3-642-13232-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13232-2.

Rutkowski:2010:AISa

[RST^{+}10b] Leszek Rutkowski, Rafał Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M. Zurada, editors. Artificial Intelligence and Soft Computing: 10th International Conference, ICAISC 2010, Zakopane, Poland, June 13–17, 2010, Part I, volume 6113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13207-3 (print), 3-642-13208-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13208-7.

Rutkowski:2004:AIS

[RSTZ04] Leszek Rutkowski, Jörg Siekmann, Ryszard Tadeusiewicz, and Lotfi A. Zadeh, editors. Artificial Intelligence and Soft Computing— ICAISC 2004: 7th International Conference, Zakopane, Poland, June 7–11, 2004, Proceedings, volume 3070 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22123-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I86 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3070.
htm;   http://www.springerlink.
com/content/978-3-540-
22123-4;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3070;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98109.

Rajapakse:2007:PRB

[RSV07] Jagath C. Rajapakse, Bertil Schmidt, and Gwenn Volkert, editors. Pattern Recognition in Bioinformatics: Second IAPR International Workshop, PRIB 2007, Singapore, October 1–2, 2007. Proceedings, volume 4774 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75285-4 (print), 3-540-75286-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75286-8.

Romijn:2005:IFM

[RSvdP05] Judi Romijn, Graeme Smith, and Jaco van de Pol, editors. Integrated formal methods: 5th international conference, IFM 2005, Eindhoven, The Netherlands, November 29–December 2, 2005: proceedings, volume 3771 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30492-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        30492-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3771.

Rovan:2012:MFC

[RSW12a] Branislav Rovan, Vladimiro Sassone, and Peter Widmayer, editors. Mathematical Foundations of Computer Science 2012: 37th International Symposium, MFCS 2012, Bratislava, Slovakia, August 27–31, 2012. Proceedings, volume 7464 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32588-2 (print), 3-642-32589-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32589-2.

Ryan:2012:ISP

[RSW12b] Mark D. Ryan, Ben Smyth, and Guilin Wang, editors. Information Security Practice and Experience: 8th International Conference, IS-PEC 2012, Hangzhou, China, April 9–12, 2012. Proceedings, volume 7232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29100-7 (print), 3-642-29101-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29101-2.

Revell:1995:DES

[RT95] Norman Revell and A. Min Tjor, editors. Database and expert systems applications: 6th international conference, DEXA ’95, London, United Kingdom, September 1995: proceedings, volume 978 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60303-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D32 D3629 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0978.
htm;   http://www.springerlink.
com/content/978-3-540-

                                        60303-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=978.

Reichel:2000:SAS

[RT00] Horst Reichel and Sophie Tison, editors. STACS 2000: 17th Annual Symposium on Theoretical Aspects of Computer Science, Lille, France, February 17–19, 2000: Proceedings, volume 1770 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67141-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S958 2000; QA267.A1 L43 no.1770. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1770.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67141-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1770.

Raskin:2007:FMA

[RT07] Jean-François Raskin and P. S. Thiagarajan, editors. Formal Modeling and Analysis of Timed Systems: 5th International Conference, FORMATS 2007, Salzburg, Austria, October 3–5, 2007. Proceedings, volume 4763 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75453-9 (print), 3-540-75454-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75454-1.

Rossi:2009:ADT

[RT09] Francesca Rossi and Alexis Tsoukias, editors. Algorithmic Decision Theory: First International Conference, ADT 2009, Venice, Italy, October 20–23, 2009. Proceedings, volume 5783 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04427-1 (print), 3-642-04428-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04428-1.

Rosenblum:2010:FAS

[RT10] David S. Rosenblum and Gabriele Taentzer, editors. Fundamental Approaches to Software Engineering: 13th International Conference, FASE 2010, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2010, Paphos, Cyprus, March 20–28, 2010. Proceedings, volume 6013 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12028-8 (print), 3-642-12029-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12029-9.

Raphael:2012:ABI

[RT12] Ben Raphael and Jijun Tang, editors. Algorithms in Bioinformatics: 12th International Workshop, WABI 2012, Ljubljana, Slovenia, September 10–12, 2012. Proceedings, volume 7534 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33121-1 (print), 3-642-33122-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
33122-0.

Reggio:1995:RTD

[RTA95] Gianna Reggio, Andrzej Tarlecki, and E. Astesiano, editors. Recent trends in data type specification: 10th Workshop on Specification of Abstract Data Types, joint with the 5th COMPASS Workshop, S. Margherita, Italy, May 30–June 3, 1994: Selected Papers, volume 906 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59132-X (Berlin), 0-387-59132-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A23 W67 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0906.htm;
http://www.springerlink.
com/content/978-3-540-
59132-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=906.

Reich:2002:HOS

[RTB02] Siegfried Reich, Manolis M. Tzagarakis, and Paul M. E. De Bra, editors. Hypermedia: openness, structural awareness, and adaptivity: International Workshops OHS-7, SC-3, and AH-3, Aarhus, Denmark, August 14– 18, 2001: Revised Papers, volume 2266 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43293-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I59 H973 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2266.
htm;   http://www.springerlink.
com/content/978-3-540-
43293-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2266.

Riano:2012:KRH

[RtTM12] David Riaño, Annette ten Teije, and Silvia Miksch, editors. Knowledge Representation for Health-Care: AIME 2011 Workshop KR4HC 2011, Bled, Slovenia, July 2– 6, 2011. Revised Selected Papers, volume 6924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27696-2 (print), 3-642-27697-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27697-2.

Riano:2010:KRH

[RtTMP10] David Riaño, Annette ten Teije, Silvia Miksch, and Mor Peleg, editors. Knowledge Representation for Health-Care. Data, Processes and Guidelines: AIME 2009 Workshop KR4HC 2009, Verona, Italy, July 19, 2009, Revised Selected and Invited Papers, volume 5943 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11807-0 (print), 3-642-11808-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11808-1.

Riano:2011:KRH

[RtTMP11] David Riaño, Annette ten Teije, Silvia Miksch, and Mor Peleg, editors. Knowledge Representation for Health-Care: ECAI 2010 Workshop KR4HC 2010, Lisbon, Portugal, August 17, 2010, Revised Selected Papers, volume 6512 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18049-3 (print), 3-642-18050-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18050-7.

Ras:2008:MCD

[RTZ08] Zbigniew W. Raś, Shusaku Tsumoto, and Djamel Zighed, editors. Mining Complex Data: ECML/ PKDD 2007 Third International Workshop, MCD 2007, Warsaw, Poland, September 17– 21, 2007, Revised Selected Papers, volume 4944 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68415-8 (print), 3-540-68416-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68416-9.

Rutkowski:2006:AIS

[RTZŻ06] Leszek Rutkowski, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M.

Żurada, editors. Artificial Intelligence and Soft Computing — ICAISC 2006: 8th International Conference, Zakopane, Poland, June 25–29, 2006. Proceedings, volume 4029 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35748-3 (print), 3-540-35750-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35750-6.

Rutkowski:2008:AIS

[RTZZ08] Leszek Rutkowski, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M. Zurada, editors. Artificial Intelligence and Soft Computing — ICAISC 2008: 9th International Conference Zakopane, Poland, June 22–26, 2008 Proceedings, volume 5097 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69572-9 (print), 3-540-69731-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69731-2.

Rucklidge:1996:EVR

[Ruc96] William Rucklidge. Efficient visual recognition using the Hausdorff distance, volume 1173 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61993-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 178 pp. LCCN TA1634.R83 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1173.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61993-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1173.

Rueppel:1993:ACE

[Rue93] Rainer A. Rueppel, editor. Advances in cryptology — EURO-CRYPT ’92: Workshop on the Theory and Application of Cryptographic Techniques, Balatonfüred, Hungary, May 24–28, 1992: proceedings, volume 658 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56413-6 (New York), 3-540-56413-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0658.htm;
 http://www.springerlink.
com/content/978-0-387-
56413-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=658.

Ramos:2004:ADS

[RUL04] Félix F. Ramos, Herwig Unger, and Victor Larios, editors. Advanced Distributed Systems: Third International School and Symposium, ISSADS 2004, Guadalajara, Mexico, January 24–30, 2004. Revised Selected Papers, volume 3061 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22172-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3061.htm;
 http://www.springerlink.
com/content/978-3-540-
22172-2;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3061;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98139.

Rus:2000:AMS

[Rus00] Teodor Rus, editor. Algebraic methodology and software technology: 8th International Conference, AMAST 2000, Iowa City, Iowa, USA, May 20–27, 2000: Proceedings, volume 1816 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67530-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1816. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1816.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67530-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1816.

Rust:2005:OST

[Rus05]  Heinrich Rust.  Op-

erational semantics for timed systems: a nonstandard approach to uniform modeling of timed and hybrid systems, volume 3456 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25576-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25576-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3456;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b135629.

Ruthing:1998:ICM

[Rüt98] Oliver Rüthing, editor.

Interacting Code Motion Transformations: Their Impact and Their Complexity, volume 1539 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65510-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 225 pp. LCCN QA76.76.C65 R88 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1539.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        65510-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1539.

Rozenberg:1998:LES

[RV98] Grzegorz Rozenberg and Frits Vaandrager, editors. Lectures on embedded systems: European Educational Forum School on Embedded Systems, Veldhoven, The Netherlands, November 25–29, 1996, volume 1494 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65193-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42E97 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1494.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65193-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1494.

Rolim:2002:RAT

[RV02] Jose D. P. Rolim and Salil Vadhan, editors. Randomization and approximation techniques in computer science: 6th International Workshop, RANDOM 2002, Cambridge, MA, USA, September 13–15, 2002: Proceedings, volume 2483 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44147-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I648 2002; QA76 .R266 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2483.
htm;   http://www.springerlink.
com/content/978-3-540-
44147-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2483.

Rovan:2003:MFC

[RV03] Branislav Rovan and Peter Vojtáš, editors. Mathematical Foundations of Computer Science 2003: 28th International Symposium, MFCS 2003, Bratislava, Slovakia, August 25–29, 2003: Proceedings, volume 2747 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40671-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 2003. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t2747.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40671-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2747;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11836.

Roli:2005:IAP

[RV05] Fabio Roli and Sergio Vitulano, editors. Image analysis and processing

— ICIAP 2005: 13th international conference, Cagliari, Italy, September 6–8, 2005. proceedings, volume 3617 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28869-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28869-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3617.

Real:2010:RST

[RV10] Jorge Real and Tullio Vardanega, editors. Reliable Software Technology — Ada-Europe 2010: 15th Ada-Europe International Conference on Reliable Software Technologies, Valencia, Spain, June 14–18, 2010. Proceedings, volume 6106 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13549-8 (print), 3-642-13550-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13550-7.

Romanovsky:2011:RST

[RV11] Alexander Romanovsky and Tullio Vardanega, editors. Reliable Software Technologies — Ada-Europe 2011: 16th Ada-Europe International Conference on Reliable Software Technologies, Edinburgh, UK, June 20– 24, 2011. Proceedings, volume 6652 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21337-5 (print), 3-642-21338-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21338-0.

Rangarajan:2005:EMM

[RVY05] Anand Rangarajan, Baba C. Vemuri, and Alan L. (Alan L.) Yuille, editors. Energy minimization methods in computer vision and pattern recognition: 5th International Workshop, EMM-CVPR 2005, St. Augustine, FL, USA, November 9–11, 2005: proceedings, volume 3757 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30287-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .E53 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        30287-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3757.

Robinet:1986:EES

[RW86] B. Robinet and R. Wilhelm, editors. ESOP 86: European Symposium on Programming, Saarbrucken, Federal Republic of Germany, March 17–19, 1986: proceedings, volume 213 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16442-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E9761 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0213.htm;
http://www.springerlink.
com/content/978-0-387-
16442-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=213.

Rensink:2006:MDA

[RW06] Arend Rensink and Jos Warmer, editors. Model Driven Architecture — Foundations and Applications: Second European Conference, ECMDA-FA 2006, Bilbao, Spain, July 10– 13, 2006. Proceedings, volume 4066 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35909-5 (print), 3-540-35910-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35910-4.

Ram:2011:CBR

[RW11] Ashwin Ram and Nirmalie Wiratunga, editors. Case-Based Reasoning Research and Development: 19th International Conference on Case-Based Reasoning, ICCBR 2011, London, UK, September 12– 15, 2011. Proceedings, volume 6880 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23290-6 (print), 3-642-23291-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23291-6.

Rajapakse:2006:PRB

[RWA06] Jagath C. Rajapakse, Limsoon Wong, and Raj Acharya, editors. Pattern Recognition in Bioinformatics: International Workshop, PRIB 2006, Hong Kong, China, August 20, 2006. Proceedings, volume 4146 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37446-9 (print), 3-540-37447-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
37447-3.

Ropo:2009:RAP

[RWD09] Matti Ropo, Jan Westerholm, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 16th European PVM/MPI Users’ Group Meeting, Espoo, Finland, September 7–10, 2009. Proceedings, volume 5759 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03769-0 (print), 3-642-03770-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03770-2.

Rahwan:2012:PPP

[RWSS12] Iyad Rahwan, Wayne Wobcke, Sandip Sen, and Toshiharu Sugawara, editors. PRIMA 2012: Principles and Practice of Multi-Agent Systems: 15th International Conference, Kuching, Sarawak, Malaysia, September 3–7, 2012. Proceedings, volume 7455 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32728-1 (print), 3-642-32729-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32729-2.

Raymond:1997:AIF

[RWY97] Darrell Ronald Raymond, Derick Wood, and Sheng Yu, editors. Automata implementation: First International Workshop on Implementing Automata, WIA’96, London, Ontario, Canada, August 29–31, 1996: revised papers, volume 1260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63174-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I5787 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1260.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63174-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1260.

Riekki:2011:AGP

[RYG11] Jukka Riekki, Mika Ylianttila, and Minyi Guo, editors. Advances in Grid and Pervasive Computing: 6th International Conference, GPC 2011, Oulu, Finland, May 11–13, 2011. Proceedings, volume 6646 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20753-7 (print), 3-642-20754-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20754-9.

Ras:1991:MIS

[RZ91] Zbigniew W. Ras and Maria Zemankova, editors. Methodologies for intelligent systems: 6th International Symposium, ISMIS ’91, Charlotte, N.C., USA, October 16–19, 1991: proceedings, volume 542 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54563-8 (Berlin), 0-387-54563-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0542.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54563-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=542.

Ras:1994:MIS

[RZ94] Zbigniew Ras and Maria Zemankova, editors. Methodologies for intelligent systems: 8th International Symposium, ISMIS ’94, Charlotte, North Carolina, USA, October 16–19, 1994: proceedings, volume 869 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58495-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 1994.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0869.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        58495-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=869.

Russo:2012:PAD

[RZ12] Claudio Russo and Neng-Fa Zhou, editors. Practical Aspects of Declarative Languages: 14th International Symposium, PADL 2012, Philadelphia, PA, USA, January 23–24, 2012. Proceedings, volume 7149 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27693-8 (print), 3-642-27694-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27694-1.

Sharma:1996:ALT

[SA96] Arun K. Sharma and S. (Setsuo) Arikawa, editors. Algorithmic learning theory: 7th international workshop, ALT ’96, Sydney, Australia, October 23–25, 1996: proceedings, volume 1160 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61863-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43A48 1996. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t1160.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61863-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1160.

Steel:1997:RAA

[SA97] Sam Steel and Rachid Alami, editors. Recent advances in AI planning: 4th European Conference on Planning, ECP ’97, Toulouse, France, September 24–26, 1997: proceedings, volume 1348 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63912-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .E37 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1348.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63912-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1348.

Suzuki:2004:DSI

[SA04] Einoshin Suzuki and S. (Setsuo) Arikawa, editors. Discovery science: 7th international conference, DS 2004, Padova, Italy, October 2–5, 2004: proceedings, volume 3245 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23357-1 (hard cover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q180.55.D57 D84 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23357-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3245;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100845.

Sichman:2006:MAB

[SA06] Jaime S. Sichman and Luís Antunes, editors. Multi-Agent-Based Simulation VI: International Workshop, MABS 2005, Utrecht, The Netherlands, July 25, 2005, Revised and Invited Papers, volume 3891 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33380-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        33380-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3891.

Sloane:2012:SLE

[SA12] Anthony Sloane and Uwe Aßmann, editors. Software Language Engineering: 4th International Conference, SLE 2011, Braga, Portugal, July 3– 4, 2011, Revised Selected Papers, volume 6940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28829-4 (print), 3-642-28830-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28830-2.

Sloot:2003:CSIa

[SAB^{+}03a] Peter M. A. Sloot, David Abramson, Alexander V. Bogdanov, Jack J. Dongarra, Albert Y. Zomaya, and Yuriy E. Gorbachev, editors. Computational science — ICCS 2003: International Conference, Melbourne, Australia and St. Petersburg, Russia, June 2– 4, 2003: Proceedings, Part I, volume 2657 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40194-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2657.htm;
 http://www.springerlink.
com/content/978-3-540-
40194-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2657.

Sloot:2003:CSIb

[SAB^{+}03b] Peter M. A. Sloot, David Abramson, Alexander V. Bogdanov, Jack J. Dongarra, Albert Y. Zomaya, and Yuriy E. Gorbachev, editors. Computational science — ICCS 2003: International Conference, Melbourne, Australia and St. Petersburg, Russia, June 2– 4, 2003: Proceedings, Part II, volume 2658 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40195-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4 2003. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2658.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40195-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2658.

Sloot:2003:CSIc

[SAB^{+}03c] Peter M. A. Sloot, David Abramson, Alexander V. Bogdanov, Jack J. Dongarra, Albert Y. Zomaya, and Yuriy E. Gorbachev, editors. Computational science — ICCS 2003: International Conference, Melbourne, Australia and St. Petersburg, Russia, June 2– 4, 2003: Proceedings, Part III, volume 2659 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40196-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2659.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40196-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2659.

Sloot:2003:CSId

[SAB^{+}03d] Peter M. A. Sloot, David Abramson, Alexander V. Bogdanov, Jack J. Dongarra, Albert Y. Zomaya, and Yuriy E. Gorbachev, editors. Computational science — ICCS 2003: International Conference, Melbourne, Australia and St. Petersburg, Russia, June 2– 4, 2003: Proceedings, Part IV, volume 2660 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40197-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2660.htm;
 http://www.springerlink.
com/content/978-3-540-
40197-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2660.

Saberi:2010:INE

[Sab10] Amin Saberi, editor.

Internet and Network Economics: 6th International Workshop, WINE 2010, Stanford, CA, USA, December 13– 17, 2010. Proceedings, volume 6484 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17571-6 (print), 3-642-17572-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17572-5.

Sachse:2004:CCM

[Sac04] Frank B. Sachse. Computational cardiology: modeling of anatomy, electrophysiology, and mechanics, volume 2966 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21907-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 322 pp. LCCN QP112.4 .S23 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        21907-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2966;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b96841.

Spaccapietra:2006:JDSa

[SAC^{+}06] Stefano Spaccapietra, Paolo Atzeni, Wesley W. Chu, Tiziana Catarci, and Katia P. Sycara, editors. Journal on Data Semantics. V, volume 3870 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31426-1. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 236 pp. LCCN QA76.7 .J68 2006. URL http:

//www.springerlink.
com/content/978-3-540-
31426-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3870.

Spaccapietra:2006:JDSb

[SACM06] Stefano Spaccapietra, Karl Aberer, and Philippe Cudré-Mauroux, editors. Journal on Data Semantics VI, volume 4090 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36712-8 (print), 3-540-36871-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36871-7.

Sadeghi:2013:FCD

[Sad13] Ahmad-Reza Sadeghi, editor. Financial Cryptography and Data Security: 17th International Conference, FC 2013 Okinawa, Japan, April 1– 5, 2013. Revised Selected Paper, volume 7859 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2013. ISBN 3-642-39884-7. xv + 406 pp. LCCN QA76.9.A25.

Spaccapietra:2007:JDS

[SAF^{+}07] Stefano Spaccapietra, Paolo Atzeni, François Fages, Mohand-Saïd Hacid, Michael Kifer, John Mylopoulos, Barbara Pernici, Pavel Shvaiko, Juan Trujillo, and Ilya Zaihrayeu, editors. Journal on Data Semantics IX, volume 4601 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74982-9 (print), 3-540-74987-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74987-5.

Sagiv:2005:PLS

[Sag05] Mooly Sagiv, editor. Programming languages and systems: 14th European Symposium on Programming, ESOP 2005, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2005, Edinburgh, UK, April 4–8, 2005: Proceedings, volume 3444 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25435-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3444;
 http://www.springerlink.
com/content/978-3-540-
25435-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3444;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b107380.

Sidorov:2010:AAI

[SAG10a] Grigori Sidorov, Arturo Hernández Aguirre, and Carlos Alberto Reyes García, editors. Advances in Artificial Intelligence: 9th Mexican International Conference on Artificial Intelligence, MICAI 2010, Pachuca, Mexico, November 8–13, 2010, Proceedings, Part I, volume 6437 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16760-8 (print), 3-642-16761-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16761-4.

Sidorov:2010:ASC

[SAG10b] Grigori Sidorov, Arturo Hernández Aguirre, and Carlos Alberto Reyes García, editors. Advances in Soft Computing: 9th Mexican International Conference on Artificial Intelligence, MICAI 2010, Pachuca, Mexico, November 8–13, 2010, Proceedings, Part II, volume 6438 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16772-1 (print), 3-642-16773-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16773-7.

Sahai:2013:TCT

[Sah13] Amit Sahai, editor. Theory of cryptography: 10th Theory of Cryptography Conference, TCC 2013, Tokyo, Japan, March 3– 6, 2013: proceedings, volume 7785 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2013. ISBN 3-642-36594-9 (paperback), 3-642-36593-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 T473 2013. URL http:/

/www.loc.gov/catdir/
enhancements/fy1310/
2013931230-d.html;
http://www.loc.gov/
catdir/enhancements/
fy1310/2013931230-t.
html;   http://www.
springerlink.com/content/
978-3-642-36594-2.

Sakata:1991:AAA

[Sak91] S. Sakata, editor. Applied algebra, algebraic algorithms, and errorcorrecting codes: 8th International Conference, AAECC-8 Tokyo, Japan, August 20–24, 1990: proceedings, volume 508 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54195-0 (Berlin), 0-387-54195-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA268 .A35 1990. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0508.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54195-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=508.

Sakellariou:2001:EPP

[Sak01] Rizos Sakellariou, editor. Euro-Par 2001 parallel processing: 7th International Euro-PAr Conference, Manchester, UK, August 28–31, 2001: Proceedings, volume 2150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42495-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I553 2001; QA267.A1 L43 no.2150. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2150.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42495-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2150.

Salwicki:1983:LPT

[Sal83] Andrzej Salwicki, editor. Logics of programs and their applications: proceedings, Poznan, August 23–29, 1980, volume 148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-11981-7, 3-540-11981-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L627 1983. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0148.htm;
 http://www.springerlink.
com/content/978-0-387-
11981-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=148.

Samelson:1976:ECP

[Sam76] Klaus Samelson, editor. ECI conference 1976: proceedings of the 1st conference of the European Cooperation in Informatics, Amsterdam, August 9–12, 1976, volume 44 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1976. CODEN LNCSD9. ISBN 0-387-07804-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .E95 1976. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t0044.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        07804-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=44.

Sanderson:1980:RTC

[San80a] John G. Sanderson. A relational theory of computing, volume 82 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-09987-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 147 pp. LCCN QA76.6 .S262. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0082.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        09987-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=82.

Sandford:1980:USM

[San80b] David M. Sandford. Using sophisticated models in resolution theorem proving, volume 90 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10231-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 239 pp. LCCN QA76.9.A96

.S3. URL http://link.
springer-ny.com/link/
service/series/0558/
tocs/t0090.htm;   http:
//www.springerlink.
com/content/978-0-387-
10231-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=90.

Revision of thesis (Ph. D.).

Sandini:1992:CVE

[San92] G. Sandini, editor. Computer vision — ECCV ’92: Second European Conference on Computer Vision, Santa Margherita Ligure, Italy, May 19–22, 1992, proceedings, volume 588 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55426-2 (Berlin), 0-387-55426-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .E87 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0588.htm;
 http://www.springerlink.
com/content/978-3-540-
55426-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=588.

Sannella:1994:PLS

[San94] D. Sannella, editor. Programming languages and systems, ESOP ’94: 5th European Symposium on Programming, Edinburgh, UK, April 1994: proceedings, volume 788 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57880-3 (USA), 3-540-57880-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0788.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57880-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=788.

Sands:2001:PLS

[San01] David Sands, editor. Programming languages and systems: 10th European Symposium on Programming, ESOP 2001, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2001, Genova, Italy, April 2– 6, 2001: Proceedings, volume 2028 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41862-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 2001; QA267.A1 L43 no.2028. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2028.htm;
 http://www.springerlink.
com/content/978-3-540-
41862-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2028.

Sander:2002:SPD

[San02] Tomas Sander, editor. Security and privacy in digital rights management: ACM CCS-8 workshop DRM 2001, Philadelphia, PA, USA, November 5, 2001: Revised Papers, volume 2320 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43677-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 A345 2001. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t2320.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43677-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2320.

Saraswat:2003:ACS

[Sar03] Vijay A. Saraswat, editor. Advances in Computing Science—ASIAN 2003: Programming Languages and Distributed Computation: 8th Asian Computing Science Conference, Mumbai, India, December 10–14, 2003: Proceedings, volume 2896 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20632-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .A735 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2896.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20632-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2896;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b94667.

Sassone:2005:FSS

[Sas05] Vladimiro Sassone, editor. Foundations of Software Science and Computational Structures: 8th International Conference, FOSSACS 2005 Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2005 Edinburgh, UK, April 4–8, 2005 Proceedings, volume 3441 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25388-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F67 2005. URL http:

//www.springerlink.
com/content/978-3-540-
25388-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3441;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106850.

Sattar:1997:ATA

[Sat97] Abdul Sattar, editor. Advanced topics in artificial intelligence: 10th Australian Joint Conference on Artificial Intelligence, AI’97, Perth, Australia, November 30– December 4, 1997, proceedings, volume 1342 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63797-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1342.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1342.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        63797-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1342.

Sombattheera:2011:MDT

[SAUL11] Chattrakul Sombattheera, Arun Agarwal, Siba K. Udgata, and Kittichai Lavangnananda, editors. Multi-disciplinary Trends in Artificial Intelligence: 5th International Workshop, MI-WAI 2011, Hyderabad, India, December 7–9, 2011. Proceedings, volume 7080 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25724-0 (print), 3-642-25725-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25725-4.

Savaci:2006:AIN

[Sav06] F. Acar Savacı, editor. Artificial Intelligence and Neural Networks: 14th Turkish Symposium, TAINN 2005, Izmir, Turkey, June 16–17, 2005, Revised Selected Papers, volume 3949 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36713-6 (print), 3-540-36861-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36861-8.

Schafer:1995:SEE

[SB95] Wilhelm Schafer and Pere Botella, editors. Software engineering — ESEC ’95: 5th European Software Engineering Conference, Sitges, Spain, September 25–28, 1995: proceedings, volume 989 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60406-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .E97 1995. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t0989.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60406-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=989.

Schneider:2007:TPH

[SB07] Klaus Schneider and Jens Brandt, editors. Theorem Proving in Higher Order Logics: 20th International Conference, TPHOLs 2007, Kaiserslautern, Germany, September 10– 13, 2007. Proceedings, volume 4732 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74590-4 (print), 3-540-74591-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74591-4.

Sirakoulis:2012:CAI

[SB12a] Georgios Ch. Sirakoulis and Stefania Bandini, editors. Cellular Automata: 10th International Conference on Cellular Automata for Research and Industry, ACRI 2012, Santorini Island, Greece, September 24–27, 2012.

Proceedings, volume 7495 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33349-4 (print), 3-642-33350-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33350-7.

Srinivasa:2012:BDA

[SB12b] Srinath Srinivasa and Vasudha Bhatnagar, editors. Big Data Analytics: First International Conference, BDA 2012, New Delhi, India, December 24–26, 2012. Proceedings, volume 7678 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35541-2 (print), 3-642-35542-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35542-4.

Shepherd:1994:NOS

[SBC^{+}94] D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors. Network and Operating System Support for Digital Audio and Video. 4th International Workshop, NOSSDAV ’93. Proceedings, Lancaster, UK, 3–5 November 1993, volume 846 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58404-8, 0-387-58404-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7881.4 .N394 1993. DM58.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0846.htm;
 http://www.springerlink.
com/content/978-3-540-
58404-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=846.

Smith:1976:MER

[SBD^{+}76] B. T. Smith, J. M. Boyle, J. J. Dongarra, B. S. Garbow, Y. Ikebe, V. C. Klema, and C. B. Moler. Matrix Eigensystem Routines: EIS-PACK Guide, volume 6 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, second edition, 1976. CODEN LNCSD9. ISBN 0-387-07546-1. ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 551 pp. LCCN QA193 .M371 1976. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0006.htm;
 http://www.springerlink.
com/content/978-0-387-
07546-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=6. See

also [GBDM77].

Sichman:2003:MAB

[SBD03] Jaime S. (Jaime Simão)

Sichman, François Bousquet, and Paul Davidsson, editors. Multiagent-based simulation II: Third International Workshop, MABS 2002, Bologna, Italy, July 15– 16, 2002: Revised Papers, volume 2581 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00607-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M32 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2581.htm;
 http://www.springerlink.
com/content/978-3-540-
00607-7;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2581.

Also available via the World Wide Web.

Stiefelhagen:2008:MTP

[SBF08] Rainer Stiefelhagen, Rachel Bowers, and Jonathan Fiscus, editors. Multimodal Technologies for Perception of Humans: International Evaluation Workshops CLEAR 2007 and RT 2007, Baltimore, MD, USA, May 8–11, 2007, Revised Selected Papers, volume 4625 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68584-7 (print), 3-540-68585-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68585-2.

Slind:2004:TPH

[SBG04] Konrad Slind, Annette Bunker, and Ganesh Gopalakrishnan, editors. Theorem proving in higher order logic: 17th international conference, TPHOLS 2004, Park City, Utah, USA, September 14–17, 2004, proceedings, volume 3223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23017-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
23017-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3223;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100400.

Sloot:1998:HPC

[SBH98] Peter Sloot, Marian Bubak, and Bob Hertzberger, editors. High-performance computing and networking: International Conference and exhibition, Amsterdam, The Netherlands, April 1998: Proceedings, volume 1401 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64443-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88.H5 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1401.
htm;   http://www.springerlink.
com/content/978-3-540-
64443-9;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1401.

Syropoulos:2004:TXD

[SBH^{+}04] Apostolos Syropoulos, Karl Berry, Yannis Haralambous, Baden Hughes, Steven Peter, and John Plaice, editors. TEX, XML, and Digital Typography: International Conference on TEX, XML, and Digital Typography, held jointly with the 25th Annual Meeting of the TEX Users Group, TUG 2004, Xanthi, Greece, August 30– September 3, 2004: Proceedings, volume 3130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3130.htm;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3130;   http:
                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99374.

Stutzle:2007:ESL

[SBH07] Thomas Stützle, Mauro Birattari, and Holger H. Hoos, editors. Engineering Stochastic Local Search Algorithms. Designing, Implementing and Analyzing Effective Heuristics, International Workshop, SLS 2007, Brussels, Belgium, September 6–8, 2007, Proceedings, volume 4638 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-74445-2. URL http:

//www.springerlink.
com/content/978-3-540-
74445-0.

Stutzle:2009:ESL

[SBH09] Thomas Stützle, Mauro Birattari, and Holger H. Hoos, editors. Engineering Stochastic Local Search Algorithms. Designing, Implementing and Analyzing Effective Heuristics: Second International Workshop, SLS 2009, Brussels, Belgium, September 3–4, 2009. Proceedings, volume 5752 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03750-X (print), 3-642-03751-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03751-1.

Spaccapietra:2005:JDS

[SBJ^{+}05] Stefano Spaccapietra, Elisa Bertino, Sushil Jajodia, Roger King, Dennis McLeod, Maria E. Orlowska, and Leon Strous, editors. Journal on Data Semantics. II, volume 3360 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24208-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 221 pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24208-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3360;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104761.

Stone:2001:RRS

[SBK01] Peter Stone, Tucker Balch, and Gerhard K. Kraetzschmar, editors. RoboCup 2000: Robot Soccer World Cup IV, volume 2019 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42185-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ210.3 .R615 2000; QA267.A1 L43 no.2019. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2019.htm;
 http://www.springerlink.
com/content/978-3-540-
42185-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2019.

Simoff:2008:VDM

[SBM08] Simeon J. Simoff, Michael H.

Böhlen, and Arturas Mazeika, editors. Visual Data Mining: Theory, Techniques and Tools for Visual Analytics, volume 4404 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-71079-5 (print), 3-540-71080-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71080-6.

Sommer:2011:RAI

[SBM11] Robin Sommer, Davide Balzarotti, and Gregor Maier, editors. Recent Advances in Intrusion Detection: 14th International Symposium, RAID 2011, Menlo Park, CA, USA, September 20– 21, 2011. Proceedings, volume 6961 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23643-X (print), 3-642-23644-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23644-0.

Sacchi:2011:MAC

[SBV^{+}11] Claudio Sacchi, Boris Bellalta, Alexey Vinel, Christian Schlegel, Fabrizio Granelli, and Yan Zhang, editors. Multiple Access Communications: 4th International Workshop, MACOM 2011, Trento, Italy, September 12–13, 2011. Proceedings, volume 6886 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23794-0 (print), 3-642-23795-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23795-9.

Srivas:1996:FMC

[SC96] Mandayam Srivas and Albert Camilleri, editors. Formal methods in computer-aided design: first international conference, FMCAD ’96, Palo Alto, CA, USA, November 6–8, 1996: proceedings, volume 1166 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61937-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7874.65.F53 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1166.
htm;   http://www.springerlink.
com/content/978-3-540-
61937-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1166.

Smyth:1998:ACB

[SC98] Barry Smyth and Padraig Cunningham, editors. Advances in case-based reasoning: 4th European workshop, EWCBR-98, Dublin, Ireland, September 23–25, 1998: Proceedings, volume 1488 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64990-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1488. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1488.htm;
 http://www.springerlink.
com/content/978-3-540-
64990-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1488.

Sventek:2000:MIA

[SC00] Joseph Sventek and G. Coulson, editors. Middleware 2000: IFIP/ ACM International Conference on Distributed Systems Platforms and Open Distributed Processing, New York, NY, USA, April 4–7, 2000: Proceedings, volume 1795 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67352-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I34 2000; QA267.A1 L43 no.1795.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1795.
htm;   http://www.springerlink.
com/content/978-3-540-
67352-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1795.

Syrotiuk:2005:AHM

[SC05] Violet R. Syrotiuk and Edgar Chávez, editors. AD-HOC, mobile, and wireless networks: 4th international conference, ADHOC-NOW 2005, Cancun, Mexico, October 6–8, 2005, proceedings, volume 3738 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29132-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29132-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3738.

Sudholt:2007:OOT

[SC07] Mario Südholt and Charles Consel, editors. Object-Oriented Technology. ECOOP 2006 Workshop Reader: ECOOP 2006 Workshops, Nantes, France, July 3–7, 2006, Final Reports, volume 4379 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71772-2 (print), 3-540-71774-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71774-4.

Serra:2009:AEP

[SC09a] Roberto Serra and Rita Cucchiara, editors. AI*IA 2009: Emergent Perspectives in Artificial Intelligence: XIth International Conference of the Italian Association for Artificial Intelligence Reggio Emilia, Italy, December 9–12, 2009 Proceedings, volume 5883 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10290-5 (print), 3-642-10291-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10291-2.

Sossai:2009:SQA

[SC09b] Claudio Sossai and Gaetano Chemello, editors. Symbolic and Quantitative Approaches to Reasoning with Uncertainty: 10th European Conference, ECSQARU 2009, Verona, Italy, July 1– 3, 2009. Proceedings, volume 5590 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02905-1 (print), 3-642-02906-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02906-6.

Scholz:2011:IAF

[SC11] Sven-Bodo Scholz and Olaf Chitil, editors. Implementation and Application of Functional Languages: 20th International Symposium, IFL 2008, Hatfield, UK, September 10–12, 2008. Revised Selected Papers, volume 5836 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24451-3 (print), 3-642-24452-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24452-0.

Stumptner:2001:AAA

[SCB01] Markus Stumptner, Dan Corbett, and Mike Brooks, editors. AI 2001: Advances in Artificial Intelligence: 14th Australian Joint Conference on Artificial Intelligence, Adelaide, Australia, December 10–14, 2001: Proceedings, volume 2256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42960-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN xii + 666. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2256.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42960-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2256.

Sion:2010:FCDb

[SCD^{+}10] Radu Sion, Reza Curtmola, Sven Dietrich, Aggelos Kiayias, Josep M. Miret, Kazue Sako, and Francesc Sebé, editors. Financial Cryptography and Data Security: FC 2010 Workshops, RLCPS, WECSR, and WLC 2010, Tenerife, Canary Islands, Spain, January 25–28, 2010, Revised Selected Papers, volume 6054 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14991-X (print), 3-642-14992-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14992-4.

Sichman:1998:MAS

[SCG98] Jaime S. Sichman, Rosaria Conte, and G. Nigel Gilbert, editors. Multiagent systems and agentbased simulation: First International Workshop, MABS ’98, Paris, France, July 4–6, 1998: Proceedings, volume 1534 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65476-3 (sc). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1534.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1534.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        65476-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1534.

Schmitz:1974:IFI

[Sch74] Paul Schmitz, editor. Internationale Fachtagung: Informationszentren in Wirtschaft und Verwaltung, Köln, 17./18. Sept. 1973. (German) [International Technical Conference: Information Centers in Economics and Administration], volume 9 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CO-DEN LNCSD9. ISBN 3-540-06703-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.9.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0009.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        06703-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=9. Spon-

sored by the Gesellschaft für Informatik, Fachausschuss 8: Methoden der Informatik für spezielle Anwendungen and the Betriebswirtschaftliches Institut für Organisation und Automation, Universität Köln. English or German.

Schoning:1986:CS

[Sch86] Uwe Schoning. Complexity and structure, volume 211 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CODEN LNCSD9. ISBN 0-387-16079-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). 99 pp. LCCN QA267 .S371 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0211.htm;
http://www.springerlink.
com/content/978-0-387-
16079-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=211.

Revised and extended version of a thesis submitted to the … University of Stuttgart, Germany, in partial fulfillment of the German ‘Habilitation’.

Schmidt:1991:MLC

[Sch91] Helmut Schmidt. Metalevel control for deductive database systems, volume 479 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-53754-6 (USA), 3-540-53754-6 (Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 155 pp. LCCN QA76.9.D3 S29 1991. DM35.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0479.htm;
 http://www.springerlink.
com/content/978-0-387-
53754-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=479.

Schulz:1992:WER

[Sch92] K. U. Schulz, editor. Word equations and related topics: 1st international workshop, IWWERT ’90, Tubingen, Germany, October 1–3, 1990: proceedings, volume 572 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55124-7 (Berlin), 0-387-55124-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 I94 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0572.htm;
 http://www.springerlink.
com/content/978-3-540-
55124-9;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=572.

Schill:1993:DOD

[Sch93a] Alexander Schill, editor.

DCE — the OSF distributed computing environment: client/server model and beyond: International DCE Workshop, Karlsruhe, Germany, October 7–8, 1993: proceedings, volume 731 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57306-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C55 I58 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0731.
htm;   http://www.springerlink.
com/content/978-3-540-
57306-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=731.

Schiper:1993:DAI

[Sch93b] Andre Schiper, editor. Distributed algorithms: 7th International Workshop, WDAG ’93, Lausanne, Switzerland, September 27–29, 1993: proceedings, volume 725 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57271-6 (Berlin), 0-387-57271-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I597 1993. DM66.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0725.
htm;   http://www.springerlink.
com/content/978-3-540-
57271-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=725.

Papers from the 7th International Workshop on Distributed Algorithms.

Schafer:1995:SPT

[Sch95] Wilhelm Schafer, editor. Software process technology: 4th European workshop, EWSPT ’95, Noordwijkerhout, The Netherlands, April 3–5, 1995: proceedings, volume 913 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59205-9 (Berlin), 0-387-59205-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 E87 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0913.htm;
 http://www.springerlink.

com/content/978-3-540-
59205-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=913.

Schal:1996:WMS

[Sch96] Thomas Schal. Workflow management systems for process organisations, volume 1096 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61401-X. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 200 pp. LCCN T58.6 .S32 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1096.htm;
 http://www.springerlink.
com/content/978-3-540-
61401-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1096.

Schlechta:1997:NLB

[Sch97a] Karl Schlechta. Nonmonotonic logics: basic concepts, results, and techniques, volume 1187 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62482-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 243 pp. LCCN Q339 .S35 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1187.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62482-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1187.

Schneider:1997:SDT

[Sch97b] Markus Schneider. Spatial data types for database systems: finite resolution geometry for geographic information systems, volume 1288 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63454-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 275 pp. LCCN G70.212.S36 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1288.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63454-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1288.

Schal:1998:WMS

[Sch98a] Thomas Schal. Workflow management systems for process organisations, volume 1096 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, second edition, 1998. CODEN LNCSD9. ISBN 3-540-65304-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 229 pp. LCCN T58.6 .S32 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1096.htm;
 http://www.springerlink.
com/content/978-3-540-
65304-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1096.

Schaub:1998:ARI

[Sch98b] Torsten Schaub. The Automation of Reasoning with Incomplete Information: From Semantic Foundations to Efficient Computation, volume 1409 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64515-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 159 pp. LCCN QA76.9.A96S42 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1409.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64515-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1409.

Scheideler:1998:URS

[Sch98c] Christian Scheideler. Universal Routing Strategies for Interconnection Networks, volume 1390 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64505-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xvii + 234 pp. LCCN QA267.A1 L43 no.1390. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1390.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64505-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1390.

Extension of the author’s Ph.D. thesis submitted to Paderborn University in 1996.

Scharstein:1999:VSU

[Sch99] Daniel Scharstein. View synthesis using stereo vision, volume 1583 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66159-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 163 pp. LCCN TA1634 .S386 1999. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1583.
htm;   http://www.springerlink.
com/content/978-3-540-
66159-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1583.

Schneier:2001:FSE

[Sch01a] Bruce Schneier, editor.

Fast software encryption: 7th International Workshop, FSE 2000, New York, NY, USA, April 10–12, 2000: Proceedings, volume 1978 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41728-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no. 1978.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1978.
htm;   http://www.springerlink.
com/content/978-3-540-
41728-6;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1978.

Schumacher:2001:OCM

[Sch01b]  Michael  Schumacher.

Objective Coordination in Multi-Agent System Engineering: Design and Implementation, volume 2039 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41982-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 149 pp. LCCN QA76.76.I58 S38 2001; QA267.A1 L43 no.2039. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2039.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41982-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2039.

Schulte:2002:PCS

[Sch02] Christian Schulte. Programming Constraint Services: High-Level Programming of Standard and New Constraint Services, volume 2302 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43371-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 176 pp. LCCN QA76.612 .S38 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2302.
htm;   http://www.springerlink.
com/content/978-3-540-
43371-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2302.

Schiper:2003:FDD

[Sch03a] Andre Schiper, editor.

Future directions in distributed computing: International Workshop, Bertinoro, Italy, June 3–7, 2002: Research and Position Papers, volume 2584 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00912-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I598 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2584.htm;
 http://www.springerlink.
com/content/978-3-540-
00912-2;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2584.

Thirty-one of the 38 papers in this volume are based on the preliminary reports that appeared in the proceedings of the 2002 International Workshop on Future Directions in Distributed Computing, held on 3–7 June 2002 at the University of Bologna Residential Center located in … Bertinoro, Italy.

Schmid:2003:ISF

[Sch03b] Ute Schmid, editor. Inductive Synthesis of Functional Programs: Universal Planning, Folding of Finite Programs, and Schema Abstraction by Analogical Reasoning, volume 2654 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40174-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xxii + 398 pp. LCCN QA76.6 .S3855 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2654.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40174-2;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2654;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12055.

Schumacher:2003:SEP

[Sch03c] Markus Schumacher, editor. Security Engineering with Patterns: origins, theoretical models, and new applications, volume 2754 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40731-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 S3536 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2754.
htm;   http://www.springerlink.
com/content/978-3-540-
40731-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2754;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11930.

Schepers:2004:SCE

[Sch04a] Henk Schepers, editor.

Software and Compilers for Embedded Systems: 8th International Workshop, SCOPES 2004, Amsterdam, The Netherlands, September 2– 3, 2004: Proceedings, volume 3199 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23035-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3199.htm;
 http://www.springerlink.
com/content/978-3-540-
23035-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3199;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99901.

Schmidt:2004:PLS

[Sch04b] David Schmidt, editor. Programming Languages and Systems: 13th European Symposium on Programming, ESOP 2004, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2004, Barcelona, Spain, March 29–April 2, 2004, Proceedings, volume 2986 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21313-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2986.htm;
 http://www.springerlink.
com/content/978-3-540-
21313-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2986;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96702.

Sloot:2004:CAI

[SCH04c] Peter M. A. Sloot, Bastien Chopard, and Alfons G. Hoekstra, editors. Cellular Automata: 6th International Conference on Cellular Automata for Research and Industry, ACRI 2004, Amsterdam, The Netherlands, October 25–28, 2004, Proceedings, volume 3305 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23596-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.C45 I58 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3305.
htm;   http://www.springerlink.

                                        com/content/978-3-540-
                                        23596-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3305;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b102055.

Schmidt:2006:RKA

[Sch06] Renate A. Schmidt, editor. Relations and Kleene Algebra in Computer Science: 9th International Conference on Relational Methods in Computer Science and 4th International Workshop on Applications of Kleene Algebra, RelMiCS/AKA 2006, Manchester, UK, August 29–September 2, 2006. Proceedings, volume 4136 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37873-1 (print), 3-540-37874-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37874-7.

Schuler:2007:OCS

[Sch07] Douglas Schuler, editor. Online Communities and Social Computing: Second International Conference, OCSC 2007, Held as Part of HCI International 2007, Beijing, China, July 22– 27, 2007. Proceedings, volume 4564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73256-X (print), 3-540-73257-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73257-0.

Sakai:2009:RSF

[SCH^{+}09a] Hiroshi Sakai, Mihir Kumar Chakraborty, Aboul Ella Hassanien, Dominik Ślezak, and William Zhu, editors. Rough Sets, Fuzzy Sets, Data Mining and Granular Computing: 12th International Conference, RSFDGrC 2009, Delhi, India, December 15–18, 2009. Proceedings, volume 5908 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10645-5 (print), 3-642-10646-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        10646-0.

Schmidt:2009:ADC

[Sch09b] Renate A. Schmidt, editor. Automated Deduction — CADE-22: 22nd International Conference on Automated Deduction, Montreal, Canada, August 2–7, 2009. Proceedings, volume 5663 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02958-2 (print), 3-642-02959-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02959-2.

Scheideler:2010:ASS

[Sch10a] Christian Scheideler, editor. Algorithms for Sensor Systems: 6th International Workshop on Algorithms for Sensor Systems, Wireless Ad Hoc Networks, and Autonomous Mobile Entities, ALGOSENSORS 2010, Bordeaux, France, July 5, 2010, Revised Selected Papers, volume 6451 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16987-2 (print), 3-642-16988-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16988-5.

Schoitsch:2010:CSR

[Sch10b] Erwin Schoitsch, editor.

Computer Safety, Reliability, and Security: 29th International Conference, SAFECOMP 2010, Vienna, Austria, September 14–17, 2010. Proceedings, volume 6351 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15650-9 (print), 3-642-15651-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15651-9.

Schmid:2011:TPT

[Sch11] Klaus Schmid, editor.

Top Productivity through Software Reuse: 12th International Conference on Software Reuse, ICSR 2011, Pohang, South Korea, June 13–17, 2011. Proceedings, volume 6727 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21346-4 (print), 3-642-21347-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21347-2.

Schmitt:2012:MME

[Sch12] Jens B. Schmitt, editor. Measurement, Modelling, and Evaluation of Computing Systems and Dependability and Fault Tolerance: 16th International GI/ITG Conference, MMB & DFT 2012, Kaiserslautern, Germany, March 19– 21, 2012. Proceedings, volume 7201 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28539-2 (print), 3-642-28540-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28540-0.

Schmidt:2007:CBS

[SCHS07] Heinz W. Schmidt, Ivica Crnković, George T. Heineman, and Judith A. Stafford, editors. Component-Based Software Engineering: 10th International Symposium, CBSE 2007, Medford, MA, USA, July 9–11, 2007. Proceedings, volume 4608 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73550-X (print), 3-540-73551-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73551-9.

Stiller:2003:GCC

[SCKR03] Burkhard Stiller, Georg Carle, Martin Karsten, and Peter Reichl, editors. Group Communications and Charges: technology and business models: 5th COST 264 International Workshop on Networked Group Communications, NGC 2003 and 3rd International Workshop on Internet Charging and QoS Technologies, ICQT 2003, Munich, Germany, September 16–19, 2003: Proceedings, volume 2816 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20051-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.887 .I58 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2816.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20051-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2816;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13249.

Schaefer:2010:PPSa

[SCKR10a] Robert Schaefer, Carlos Cotta, Joanna Kołodziej, and Günter Rudolph, editors. Parallel Problem Solving from Nature, PPSN XI: 11th International Conference, Kraków, Poland, September 11–15, 2010, Proceedings, Part I, volume 6238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15843-9 (print), 3-642-15844-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15844-5.

Schaefer:2010:PPSb

[SCKR10b] Robert Schaefer, Carlos Cotta, Joanna Kołodziej, and Günter Rudolph, editors. Parallel Problem Solving from Nature, PPSN XI: 11th International Conference, Kraków, Poland, September 11–15, 2010, Proceedings, Part II, volume 6239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15870-6 (print), 3-642-15871-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15871-1.

Shen:2006:CSC

[SCL^{+}06] Weiming Shen, Kuo-Ming Chao, Zongkai Lin, Jean-Paul A. Barthès, and Anne James, editors. Computer Supported Cooperative Work in Design II: 9th International Conference, CSCWD 2005, Coventry, UK, May 24–26, 2005, Revised Selected Papers, volume 3865 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32969-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32969-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3865.

Schmidt:1988:ADT

[SCM88] Joachim W. Schmidt, Stefano Ceri, and M. Missikoff, editors. Advances in database technology, EDBT ’88: International Conference on Extending Database Technology, Venice, Italy, March 14–18, 1988, proceedings, volume 303 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-19074-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I558441 1988.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0303.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        19074-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=303.

Scott:2002:AIM

[Sco02] Donia R. Scott, editor. Artificial Intelligence: Methodology, Systems, and Applications: 10th International Conference, AIMSA 2002, Varna, Bulgaria, September 4– 6, 2002. Proceedings, volume 2443 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44127-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I554 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2443.htm;
 http://www.springerlink.
com/content/978-3-540-
44127-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2443.

Simperl:2012:SWR

[SCP^{+}12] Elena Simperl, Philipp Cimiano, Axel Polleres, Oscar Corcho, and Valentina Presutti, editors. The Semantic Web: Research and Applications: 9th Extended Semantic Web Conference, ESWC 2012, Heraklion, Crete, Greece, May 27– 31, 2012. Proceedings, volume 7295 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30283-1 (print), 3-642-30284-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30284-8.

Sichman:2006:AAI

[SCR06] Jaime Simão Sichman, Helder Coelho, and Solange Oliveira Rezende, editors. Advances in Artificial Intelligence — IBERAMIA-SBIA 2006: 2nd International Joint Conference, 10th Ibero– American Conference on AI, 18th Brazilian AI Symposium, Ribeirão Preto, Brazil, October 23–27, 2006. Proceedings, volume 4140 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45462-4 (print), 3-540-45464-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        45464-9.

Salerno:2012:CIM

[SÇS12] Emanuele Salerno, A. Enis Çetin, and Ovidio Salvetti, editors. Computational Intelligence for Multimedia Understanding: International Workshop, MUSCLE 2011, Pisa, Italy, December 13–15, 2011, Revised Selected Papers, volume 7252 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32435-5 (print), 3-642-32436-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32436-9.

Sole-Casals:2010:ANS

[SCZ10] Jordi Solé-Casals and Vladimir Zaiats, editors. Advances in Nonlinear Speech Processing: International Conference on Nonlinear Speech Processing, NOLISP 2009, Vic, Spain, June 25– 27. Revised Selected Papers, volume 5933 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11508-X (print), 3-642-11509-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11509-7.

Sangiorgi:1998:CCT

[Sd98] Davide Sangiorgi and Robert de Simone, editors. CONCUR’98: concurrency theory: 9th International Conference, Nice, France, September 8–11, 1998: Proceedings, volume 1466 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64896-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1466.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64896-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1466.

Sure:2006:SWR

[SD06] York Sure and John Domingue, editors. The Semantic Web: Research and Applications, volume 4011 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34544-2 (softcover). ???? pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34544-2.

Spaccapietra:2009:JDSc

[SD09] Stefano Spaccapietra and Lois Delcambre, editors. Journal on Data Semantics XIV, volume 5880 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10561-0 (print), 3-642-10562-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10562-3.

Stiller:2010:MAM

[SD10] Burkhard Stiller and Filip De Turck, editors. Mechanisms for Autonomous Management of Networks and Services: 4th International Conference on Autonomous Infrastructure, Management and Security, AIMS 2010, Zurich, Switzerland, June 23– 25, 2010. Proceedings, volume 6155 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13985-X (print), 3-642-13986-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13986-4.

Sherwani:1991:CFG

[SdDK91] N. A. Sherwani, E. de Doncker, and J. A. Kapenga, editors. Computing in the 90’s: the First Great Lakes Computer Science Conference, Kalamazoo, Michigan, USA, October 18–20, 1989: proceedings, volume 507 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-97628-0 (U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .G74 1989. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0507.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        97628-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=507.

Schiele:2007:AIE

[SDG^{+}07] Bernt Schiele, Anind K.

Dey, Hans Gellersen, Boris de Ruyter, Manfred Tscheligi, Reiner Wichert, Emile Aarts, and Alejandro Buchmann, editors. Ambient Intelligence: European Conference, AmI 2007, Darmstadt, Germany, November 7–10, 2007. Proceedings, volume 4794 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76651-0 (print), 3-540-76652-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76652-0.

Stefani:2003:DAI

[SDH03] Jean-Bernard Stefani, Isabelle Demeure, and Daniel Hagimont, editors. Distributed Applications and Interoperable Systems: 4th IFIP WG6.1 International Conference, DAIS 2003, Paris France November 17–21, 2003: Proceedings, volume 2893 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20529-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I33843 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2893.
htm;   http://www.springerlink.
com/content/978-3-540-
20529-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2893;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94128.

Shan:2005:TSI

[SDH05] Ming-Chien Shan, Umeshwar Dayal, and Meichun Hsu, editors. Technologies for e-services: 5th international workshop, TES 2004, Toronto, Canada, August 29–30, 2004: Revised selected papers, volume 3324 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25049-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88813 .T47 2004. URL http:

//www.springerlink.
com/content/978-3-540-
25049-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

0302-9743&volume=3324;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106626.

Stenstrom:2008:HPE

[SDK^{+}08] Per Stenström, Michel Dubois, Manolis Katevenis, Rajiv Gupta, and Theo Ungerer, editors. High Performance Embedded Architectures and Compilers: Third International Conference, HiPEAC 2008, Göteborg, Sweden, January 27– 29, 2008. Proceedings, volume 4917 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77559-5 (print), 3-540-77560-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77560-7.

Stumme:2005:CSC

[SDM05] Gerd Stumme, Frithjof Dau, and Marie-Laure Mugnier, editors. Conceptual structures: common semantics for sharing knowledge: 13th International Conference on Conceptual Structures, ICCS 2005, Kassel, Germany, July 17– 22, 2005: proceedings, volume 3596 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27783-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
27783-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3596.

Sommer:1997:CAI

[SDP97] Gerald Sommer, Kostas Daniilidis, and Josef Pauli, editors. Computer analysis of images and patterns: 7th International Conferencce, CAIP’97: Kiel, Germany, September 1997: proceedings, volume 1296 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63460-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637.I54 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1296.
htm;   http://www.springerlink.

com/content/978-3-540-
63460-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1296.

Schoenauer:2000:PPS

[SDR^{+}00] Marc Schoenauer, K. Deb, G. Rudolph, X. Yao, E. Lutton, J. J. Merelo, and H.-P. Schwefel, editors. Parallel problem solving from nature — PPSN VI: 6th International Conference, Paris, France, September 2000: Proceedings, volume 1917 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41056-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1917. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1917.htm;
http://www.springerlink.
com/content/978-3-540-
41056-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1917.

Salah:2012:HBU

[SdSMO12] Albert Ali Salah, Javier Ruiz del Solar, Çetin Meriçli, and Pierre-Yves Oudeyer, editors. Human Behavior Understanding: Third International Workshop, HBU 2012, Vilamoura, Portugal, October 7, 2012. Proceedings, volume 7559 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34013-X (print), 3-642-34014-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34014-7.

Schneider:1994:GTC

[SE94] Hans-Jurgen Schneider and Hartmut Ehrig, editors. Graph transformations in computer science: International Workshop, Dagstuhl Castle, Germany, January 4–8, 1993: proceedings, volume 776 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57787-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .G725 1994. DM80.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0776.htm;
 http://www.springerlink.
com/content/978-0-387-
57787-6;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=776.

Schwenker:2010:ANN

[SE10]  Friedhelm  Schwenker

and Neamat El Gayar, editors. Artificial Neural Networks in Pattern Recognition: 4th IAPR TC3 Workshop, ANNPR 2010, Cairo, Egypt, April 11–13, 2010. Proceedings, volume 5998 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12158-6 (print), 3-642-12159-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12159-3.

Sebastiani:2003:AIR

[Seb03] Fabrizio Sebastiani, editor. Advances in information retrieval: 25th European Conference on IR Research, ECIR 2003, Pisa, Italy, April 14– 16, 2003: Proceedings, volume 2633 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-01274-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B429 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2633.htm;
 http://www.springerlink.
com/content/978-3-540-
01274-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2633.

Also available via the World Wide Web.

Schmorrow:2009:FAC

[SEG09] Dylan D. Schmorrow, Ivy V. Estabrooke, and Marc Grootjen, editors. Foundations of Augmented Cognition. Neuroergonomics and Operational Neuroscience: 5th International Conference, FAC 2009 Held as Part of HCI International 2009 San Diego, CA, USA, July 19–24, 2009 Proceedings, volume 5638 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02811-X (print), 3-642-02812-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02812-0.

Sehr:1997:LCP

[Seh97] David Sehr, editor. Languages and compilers for parallel computing: 9th International Workshop, LCPC’96, San Jose, California, USA, August 8– 10, 1996: proceedings, volume 1239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63091-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1239.htm;
http://www.springerlink.
com/content/978-3-540-
63091-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1239.

Seidl:2007:FSS

[Sei07] Helmut Seidl, editor.

Foundations of Software Science and Computational Structures: 10th International Conference, FOSSACS 2007, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2007, Braga, Portugal, March 24–April 1, 2007. Proceedings, volume 4423 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71388-3 (print), 3-540-71389-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71389-0.

Seidl:2012:PLS

[Sei12] Helmut Seidl, editor. Programming Languages and Systems: 21st European Symposium on Programming, ESOP 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7211 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28868-5 (print), 3-642-28869-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28869-2.

Shi:2011:TDH

[SEK^{+}11] Yun Q. Shi, Sabu Emmanuel, Mohan S.

Kankanhalli, Shih-Fu Chang, Regunathan Radhakrishnan, Fulong Ma, and Li Zhao, editors. Transactions on Data Hiding and Multimedia Security VI, volume 6730 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24555-2 (print), 3-642-24556-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24556-5.

Selman:1986:SCT

[Sel86] Alan L. Selman, editor. Structure in complexity theory: proceedings of the conference held at the University of California, Berkeley, California, June 2– 5, 1986, volume 223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16486-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.223. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0223.htm;
 http://www.springerlink.

                                        com/content/978-0-387-
                                        16486-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=223.

Papers presented at the Structure in Complexity Theory Conference.

Sellis:1995:RDS

[Sel95] Timos Sellis, editor. Rules in database systems: second international workshop, RIDS ’95, Glyfada, Athens, Greece, September 25– 27, 1995: proceedings, volume 985 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60365-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I594 1995.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0985.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60365-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=985.

Sips:2009:EPP

[SEL09] Henk Sips, Dick Epema, and Hai-Xiang Lin, editors. Euro-Par 2009 Parallel Processing: 15th International Euro-Par Conference, Delft, The Netherlands, August 25– 28, 2009. Proceedings, volume 5704 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03868-9 (print), 3-642-03869-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03869-3.

Song:2007:DWK

[SEN07] Il Yeal Song, Johann Eder, and Tho Manh Nguyen, editors. Data Warehousing and Knowledge Discovery: 9th International Conference, DaWaK 2007, Regensburg Germany, September 3–7, 2007. Proceedings, volume 4654 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74552-1 (print), 3-540-74553-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74553-2.

Song:2008:DWK

[SEN08] Il-Yeol Song, Johann Eder, and Tho Manh Nguyen, editors. Data Warehousing and Knowledge Discovery: 10th International Conference, DaWaK 2008 Turin, Italy, September 2–5, 2008 Proceedings, volume 5182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85835-0 (print), 3-540-85836-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85836-2.

Sendrier:2010:PQC

[Sen10] Nicolas Sendrier, editor. Post-Quantum Cryptography: Third International Workshop, PQCrypto 2010, Darmstadt, Germany, May 25–28, 2010. Proceedings, volume 6061 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-12928-5. LCCN QA76.9.A25 2010. URL

                                        http://www.springerlink.
                                        com/content/978-3-642-
                                        12928-5.

Seznec:2009:HPE

[SEO^{+}09] André Seznec, Joel Emer, Michael O’Boyle, Margaret Martonosi, and Theo Ungerer, editors. High Performance Embedded Architectures and Compilers: Fourth International Conference, HiPEAC 2009, Paphos, Cyprus, January 25– 28, 2009. Proceedings, volume 5409 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-92989-4 (print), 3-540-92990-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92990-1.

Serna:2012:AGT

[Ser12] Maria Serna, editor. Algorithmic game theory: 5th International Symposium, SAGT 2012, Barcelona, Spain, October 22–23, 2012. Proceedings, volume 7615 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33995-6, 3-642-33996-4 (e-book). LCCN QA269 .S24 2012. URL http:

//www.springerlink.
com/content/978-3-642-
33995-0;   http://www.

                                        springerlink.com/content/
                                        978-3-642-33996-7.

Sestoft:2006:PLS

[Ses06] Peter Sestoft, editor. Programming Languages and Systems: 15th European Symposium on Programming, ESOP 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 27–28, 2006. Proceedings, volume 3924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33095-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33095-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3924.

Sexton:2009:DFF

[Sex09] Alan P. Sexton, editor. Dataspace: The Final Frontier: 26th British National Conference on Databases, BNCOD 26, Birmingham, UK, July 7–9, 2009. Proceedings, volume 5588 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02842-X (print), 3-642-02843-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02843-4.

Schlender:1974:DFP

[SF74] Bodo Schlender and Wolfgang Frielinghaus, editors. 3. [i.e., drei] Fachtagung über Programmiersprachen, Kiel, 5.–7. Marz 1974. (German) [Third Technical Conference on Programming Languages, Kiel, 5– 7 March, 1974], volume 7 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CO-DEN LNCSD9. ISBN 3-540-06666-7 (Berlin) 0-387-06666-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .G47a.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0007.
htm;   http://www.springerlink.
com/content/978-3-540-
06666-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=7. Third

special meeting of the Gesellschaft für Informatik. On spine: Proceedings 1974. German and English.

Smith:1996:ACB

[SF96] Ian Smith and Boi Faltings, editors. Advances in case-based reasoning: Third European Workshop, EWCBR-96, Lausanne, Switzerland, November 14–16, 1996: proceedings, volume 1168 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61955-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95E935 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1168.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61955-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1168.

Schrijvers:2008:CHR

[SF08a] Tom Schrijvers and Thom Frühwirth, editors. Constraint Handling Rules: Current Research Topics, volume 5388 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92242-3 (print), 3-540-92243-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92243-8.

Shvartsman:2008:SIC

[SF08b] Alexander A. Shvartsman and Pascal Felber, editors. Structural Information and Communication Complexity: 15th International Colloquium, SIROCCO 2008 Villars-sur-Ollon, Switzerland, June 17– 20, 2008 Proceedings, volume 5058 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69326-2 (print), 3-540-69355-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69355-0.

Srihari:2008:CFS

[SF08c] Sargur N. Srihari and Katrin Franke, editors. Computational Forensics: Second International Workshop, IWCF 2008, Washington, DC, USA, August 7–8, 2008. Proceedings, volume 5158 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85302-2 (print), 3-540-85303-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85303-9.

Siddiqui:2010:GRM

[SF10] Mumtaz Siddiqui and Thomas Fahringer, editors. Grid Resource Management: On-demand Provisioning, Advance Reservation, and Capacity Planning of Grid Resources, volume 5951 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CO-DEN LNCSD9. ISBN 3-642-11578-0 (print), 3-642-11579-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 206 (est.) pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11579-0.

Schmorrow:2011:FAC

[SF11] Dylan D. Schmorrow and Cali M. Fidopiastis, editors. Foundations of Augmented Cognition. Directing the Future of Adaptive Systems: 6th International Conference, FAC 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011. Proceedings, volume 6780 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21851-2 (print), 3-642-21852-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21852-1.

Silva:2011:GPE

[SFN^{+}11] Sara Silva, James A.

Foster, Miguel Nicolau, Penousal Machado, and Mario Giacobini, editors. Genetic Programming: 14th European Conference, EuroGP 2011, Torino, Italy, April 27– 29, 2011. Proceedings, volume 6621 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20406-6 (print), 3-642-20407-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20407-4.

Sako:2011:CFI

[SFS11] Hiroshi Sako, Katrin Y. Franke, and Shuji Saitoh, editors. Computational Forensics: 4th International Workshop, IWCF 2010, Tokyo, Japan, November 11–12, 2010, Revised Selected Papers, volume 6540 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19375-7 (print), 3-642-19376-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19376-7.

Strang:2011:CTV

[SFV^{+}11] Thomas Strang, Andreas Festag, Alexey Vinel, Rashid Mehmood, Cristina Rico Garcia, and Matthias Röckl, editors. Communication Technologies for Vehicles: Third International Workshop, Nets4Cars/Nets4Trains 2011, Oberpfaffenhofen, Germany, March 23– 24, 2011. Proceedings, volume 6596 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19785-X (print), 3-642-19786-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19786-4.

Stylianou:2007:PNS

[SFZE07] Yannis Stylianou, Marcos Faundez-Zanuy, and Anna Esposito, editors. Progress in Nonlinear Speech Processing, volume 4391 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71503-7 (print), 3-540-71505-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71505-4.

SparckJones:1996:ENL

[SG96] Karen Sparck Jones and J. R. (Julia Rose) Galliers. Evaluating natural language processing systems: an analysis and review, volume 1083 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61309-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). various pp. LCCN QA76.9.N38S74 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1083.
htm;   http://www.springerlink.
com/content/978-3-540-
61309-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1083.

Sun:2001:SLP

[SG01] Ron Sun and C. Lee Giles, editors. Sequence Learning: Paradigms, Algorithms, and Applications, volume 1828 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41597-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 387 pp. LCCN Q325.5 .S47 2001. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1828.
htm;   http://www.springerlink.
com/content/978-3-540-
41597-8;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1828.

Snekkenes:2003:CSE

[SG03] Einar Snekkenes and Dieter Gollmann, editors. Computer Security: ES-ORICS 2003: 8th European Symposium on Research in Computer Security, Gjøvik, Norway, October 13–15, 2003: Proceedings, volume 2808 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20300-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E98 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2808.
htm;   http://www.springerlink.
com/content/978-3-540-
20300-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2808;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13237.

Stiefelhagen:2007:MTP

[SG07] Rainer Stiefelhagen and John Garofolo, editors. Multimodal Technologies for Perception of Humans: First International Evaluation Workshop on Classification of Events, Activities and Relationships, CLEAR 2006, Southampton, UK, April 6–7, 2006, Revised Selected Papers, volume 4122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69567-2 (print), 3-540-69568-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69568-4.

Setola:2009:CII

[SG09] Roberto Setola and Stefan Geretshuber, editors. Critical Information Infrastructure Security: Third International Workshop, CRITIS 2008, Rome, Italy, October13-15, 2008. Revised Papers, volume 5508 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03551-5 (print), 3-642-03552-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
03552-4.

Sempere:2010:GIT

[SG10] José M. Sempere and Pedro García, editors. Grammatical Inference: Theoretical Results and Applications: 10th International Colloquium, ICGI 2010, Valencia, Spain, September 13– 16, 2010. Proceedings, volume 6339 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15487-5 (print), 3-642-15488-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15488-1.

Strassner:2009:MAC

[SGD09] John C. Strassner and Yacine M. Ghamri-Doudane, editors. Modelling Autonomic Communications Environments: Fourth IEEE International Workshop, MACE 2009, Venice, Italy, October 26–27, 2009. Proceedings, volume 5844 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05005-0 (print), 3-642-05006-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05006-0.

Saunders:2006:SLS

[SGG06] Craig Saunders, Marko Grobelnik, and Steve Gunn, editors. Subspace, Latent Structure and Feature Selection: Statistical and Optimization Perspectives Workshop, SLSPS 2005, Bohinj, Slovenia, February 23–25, 2005, Revised Selected Papers, volume 3940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34137-4 (softcover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34137-6.

Salakoski:2006:ANL

[SGPP06] Tapio Salakoski, Filip Ginter, Sampo Pyysalo, and Tapio Pahikkala, editors. Advances in Natural Language Processing: 5th International Conference on NLP, Fin-TAL 2006 Turku, Finland, August 23–25, 2006 Proceedings, volume 4139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37334-9 (print), 3-540-37336-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37336-0.

Salah:2010:HBU

[SGSV10] Albert Ali Salah, Theo Gevers, Nicu Sebe, and Alessandro Vinciarelli, editors. Human Behavior Understanding: First International Workshop, HBU 2010, Istanbul, Turkey, August 22, 2010. Proceedings, volume 6219 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14714-3 (print), 3-642-14715-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14715-9.

Schroeder-Heister:1991:ELP

[SH91] Peter Joseph Schroeder-Heister, editor. Extensions of logic programming: international workshop, Tubingen, FRG, December 8–10, 1989: proceedings, volume 475 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-53590-X (Berlin):, 0-387-53590-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .E97 1991. DM52.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0475.
htm;   http://www.springerlink.
com/content/978-3-540-
53590-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=475.

Speed:2007:RCM

[SH07] Terry Speed and Haiyan Huang, editors. Research in Computational Molecular Biology: 11th Annual International Conference, RECOMB 2007, Oakland, CA, USA, April 21–25, 2007. Proceedings, volume 4453 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71680-7 (print), 3-540-71681-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71681-5.

Schieferdecker:2008:MDA

[SH08] Ina Schieferdecker and Alan Hartman, editors. Model Driven Architecture — Foundations and Applications: 4th European Conference, ECMDA-FA 2008, Berlin, Germany, June 9–13, 2008. Proceedings, volume 5095 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69095-6 (print), 3-540-69100-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69100-6.

Spyropoulos:2009:SOS

[SH09] Thrasyvoulos Spyropoulos and Karin Anna Hummel, editors. Self-Organizing Systems: 4th IFIP TC 6 International Workshop, IWSOS 2009, Zurich, Switzerland, December 9–11, 2009. Proceedings, volume 5918 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10864-4 (print), 3-642-10865-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10865-5.

Steinfeld:2010:ISP

[SH10] Ron Steinfeld and Philip Hawkes, editors. Information Security and Privacy: 15th Australasian Conference, ACISP 2010, Sydney, Australia, July 5–7, 2010. Proceedings, volume 6168 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14080-7 (print), 3-642-14081-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14081-5.

Szekely:2011:IPM

[SH11] Gábor Székely and Horst K. Hahn, editors. Information Processing in Medical Imaging: 22nd International Conference, IPMI 2011, Kloster Irsee, Germany, July 3–8, 2011. Proceedings, volume 6801 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22091-6 (print), 3-642-22092-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22092-0.

Sanner:2012:RAR

[SH12a] Scott Sanner and Marcus Hutter, editors. Recent Advances in Reinforcement Learning: 9th European Workshop, EWRL 2011, Athens, Greece, September 9–11, 2011, Revised Selected Papers, volume 7188 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29945-8 (print), 3-642-29946-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29946-9.

Schindler:2012:CSC

[SH12b] Werner Schindler and Sorin A. Huss, editors. Constructive Side-Channel Analysis and Secure Design: Third International Workshop, COSADE 2012, Darmstadt, Germany, May 3–4, 2012. Proceedings, volume 7275 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29911-3 (print), 3-642-29912-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29912-4.

Shapiro:1986:TIC

[Sha86] Ehud Y. Shapiro, editor. Third International Conference on Logic Programming, Imperial College of Science and Technology, London, United Kingdom: proceedings, volume 225 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16492-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I54661 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0225.htm;
 http://www.springerlink.
com/content/978-0-387-
16492-2;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=225.

Sharp:2004:HLH

[Sha04] Richard (Richard William)

Sharp. Higher Level Hardware Synthesis, volume 2963 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21306-6. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 195 pp. LCCN TK7885.7 .S52 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2963.htm;
http://www.springerlink.
com/content/978-3-540-
21306-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2963;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95732.

Shao:2007:PLS

[Sha07] Zhong Shao, editor. Programming Languages and Systems: 5th Asian Symposium, APLAS 2007, Singapore, November 29– December 1, 2007. Proceedings, volume 4807 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76636-7 (print), 3-540-76637-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76637-7.

Sherman:1985:PLU

[She85] Mark Steven Sherman. Paragon: a language using type hierarchies for the specification, implementation, and selection of abstract data types, volume 189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CODEN LNCSD9. ISBN 0-387-15212-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 364 pp. LCCN QA76.73.P19 S54 1985.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0189.
htm;   http://www.springerlink.
com/content/978-0-387-
15212-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=189.

Based on the author’s thesis (doctoral) — Carnegie-Mellon University, 1983.

Shepherd:2001:IDM

[She01] W. D. Shepherd, editor. Interactive distributed multimedia systems: 8th International Workshop, IDMS 2001, Lancaster, UK, September 4–7, 2001: Proceedings, volume 2158 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42530-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2158; QA76.76.I59 I586 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2158.htm;
http://www.springerlink.
com/content/978-3-540-
42530-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2158.

Sherratt:2003:TBB

[She03] Edel Sherratt, editor.

Telecommunications and beyond: the broader applicability of SDL and MSC: Third International Workshop, SAM 2002, Aberystwyth , UK, June 24– 26, 2002: Revised Papers, volume 2599 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00877-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.A1 S36 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2599.htm;
 http://www.springerlink.
com/content/978-3-540-
00877-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2599.

Seipel:2005:ADP

[SHGB05] Dietmar Seipel, Michael Hanus, Ulrich Geske, and Oskar Bartenstein, editors. Applications of declarative programming and knowledge management: 15th International Conference on Applications of Declarative Programming and Knowledge Management, INAP 2004, and 18th Workshop on Logic Programming, WLP 2004, Potsdam, Germany, March 4–6, 2004. Revised selected papers, volume 3392 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25560-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.615 .I57 2004. URL http:

//www.springerlink.
com/content/978-3-540-
25560-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3392;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b135595.

Shi:2006:TDH

[Shi06] Yun Q. Shi, editor.

Transactions on Data Hiding and Multimedia Security I, volume 4300 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49071-X (print), 3-540-49072-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49072-2.

Shi:2007:TDH

[Shi07] Yun Q. Shi, editor.

Transactions on Data Hiding and Multimedia Security II, volume 4499 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CO-DEN LNCSD9. ISBN 3-540-73091-5 (print), 3-540-73092-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .T7283 2008eb. URL http:

//www.springerlink.
com/content/978-3-540-
73092-7.

Shi:2009:TDH

[Shi09] Yun Q. Shi, editor. Transactions on Data Hiding and Multimedia Security IV, volume 5510 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01756-8 (print), 3-642-01757-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01757-5.

Shi:2010:TDH

[Shi10] Yun Q. Shi, editor. Transactions on Data Hiding and Multimedia Security V, volume 6010 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14297-4 (print), 3-642-14298-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14298-7.

Shi:2012:TDHa

[Shi12] Yun Q. Shi, editor.

Transactions on Data Hiding and Multimedia Security VII, volume 7110 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28692-5 (print), 3-642-28693-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28693-3.

Salberg:2009:IAS

[SHJ09] Arnt-Børre Salberg, Jon Yngve Hardeberg, and Robert Jenssen, editors. Image Analysis: 16th Scandinavian Conference, SCIA 2009, Oslo, Norway, June 15– 18, 2009. Proceedings, volume 5575 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02229-4 (print), 3-642-02230-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02230-2.

Sattar:2006:AAA

[ShK06] Abdul Sattar and Byeong ho Kang, editors. AI 2006: Advances in Artificial Intelligence: 19th Australian Joint Conference on Artificial Intelligence, Hobart, Australia, December 4–8, 2006. Proceedings, volume 4304 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49787-0 (print), 3-540-49788-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49788-2.

Sojka:2008:TSD

[SHKP08] Petr Sojka, Aleš Horák, Ivan Kopeček, and Karel Pala, editors. Text, Speech and Dialogue: 11th International Conference, TSD 2008, Brno, Czech Republic, September 8–12, 2008. Proceedings, volume 5246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87390-2 (print), 3-540-87391-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87391-4.

Sojka:2010:TSD

[SHKP10] Petr Sojka, Aleš Horák, Ivan Kopeček, and Karel Pala, editors. Text, Speech and Dialogue: 13th International Conference, TSD 2010, Brno, Czech Republic, September 6–10, 2010. Proceedings, volume 6231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15759-9 (print), 3-642-15760-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15760-8.

Sojka:2012:TSD

[SHKP12] Petr Sojka, Aleš Horák, Ivan Kopeček, and Karel Pala, editors. Text, Speech and Dialogue: 15th International Conference, TSD 2012, Brno, Czech Republic, September 3–7, 2012. Proceedings, volume 7499 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32789-3 (print), 3-642-32790-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32790-2.

Stapleton:2008:DRI

[SHL08] Gem Stapleton, John Howse, and John Lee, editors. Diagrammatic Representation and Inference: 5th International Conference, Diagrams 2008, Herrsching, Germany, September 19– 21, 2008. Proceedings, volume 5223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87729-0 (print), 3-540-87730-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87730-1.

Sato:2010:BLL

[SHM^{+}10] Mitsuhisa Sato, Toshihiro Hanawa, Matthias S. Müller, Barbara M. Chapman, and Bronis R. de Supinski, editors. Beyond Loop Level Parallelism in OpenMP: Accelerators, Tasking and More: 6th International Workshop on OpenMP, IWOMP 2010, Tsukuba, Japan, June 14–16, 2010 Proceedings, volume 6132 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13216-2 (print), 3-642-13217-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13217-9.

Sakurai:2007:NFA

[SHN07] Akito Sakurai, Kôiti Hasida, and Katsumi Nitta, editors. New Frontiers in Artificial Intelligence: JSAI 2003 and JSAI 2004 Conferences and Workshops, Niigata, Japan, June 23– 27, 2003 and Kanazawa, Japan, May 31 — June 4, 2004, Revised Selected Papers, volume 3609 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71008-6 (print), 3-540-71009-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71009-7.

Shostak:1984:ICA

[Sho84] Robert Shostak, editor. 7th International Conference on Automated Deduction, Napa, California, USA, May 14–16, 1984: proceedings, volume 170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-96022-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 C66 1984. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0170.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        96022-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=170.

Swierstra:1999:AFP

[SHO99] S. Doaitse Swierstra, Pedro R. Henriques, and Jose N. Oliveira, editors. Advanced functional programming: Third International School, AFP ’98, Braga, Portugal, September 12–19, 1998: revised lectures, volume 1608 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66241-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1608. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1608.htm;
 http://www.springerlink.
com/content/978-3-540-
66241-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1608.

Shoup:2005:ACC

[Sho05] Victor Shoup, editor.

Advances in cryptology: CRYPTO 2005: 25th Annual International Cryptology Conference, Santa Barbara, California, USA, August 14–18, 2005: proceedings, volume 3621 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28114-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 2005; QA76.9 .A25; QA76.9 C79 2005; QA76.9 C794 2005eb; QA76.9; Internet. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28114-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3621.

Scharfe:2006:CSI

[SHØ06] Henrik Schärfe, Pascal Hitzler, and Peter Øhrstrøm, editors. Conceptual Structures: Inspiration and Application: 14th International Conference on Conceptual Structures, ICCS 2006, Aalborg, Denmark, July 16–21, 2006. Proceedings, volume 4068 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35893-5 (print), 3-540-35902-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        35902-9.

Sloot:2005:AGC

[SHP^{+}05] Peter M. A. Sloot, Alfons G. Hoekstra, Thierry Priol, Alexander Reinefeld, and Marian Bubak, editors. Advances in Grid Computing — EGC 2005: European Grid Conference, Amsterdam, The Netherlands, February 14–16, 2005, Revised Selected Papers, volume 3470 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26918-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 E974 2005.

URL http://www.springerlink.
com/content/978-3-540-
26918-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3470.

Sugimoto:2006:DLA

[SHRM06] Shigeo Sugimoto, Jane Hunter, Andreas Rauber, and Atsuyuki Morishima, editors. Digital Libraries: Achievements, Challenges and Opportunities: 9th International Conference on Asian Digital Libraries, ICADL 2006, Kyoto, Japan, November 27–30, 2006. Proceedings, volume 4312 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49375-1 (print), 3-540-49377-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
49377-8.

Szczuka:2007:AHI

[SHŚ^{+}07] Marcin S. Szczuka, Daniel Howard, Dominik Ślȩzak, Haeng kon Kim, Tai hoon Kim, Il seok Ko, Geuk Lee, and Peter M. A. Sloot, editors. Advances in Hybrid Information Technology: First International Conference, ICHIT 2006, Jeju Island, Korea, November 9–11, 2006, Revised Selected Papers, volume 4413 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77367-3 (print), 3-540-77368-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77368-9.

Stiller:2010:IOE

[SHS10] Burkhard Stiller, Tobias Hoßfeld, and George D. Stamoulis, editors. Incentives, Overlays, and Economic Traffic Control: Third International Workshop, ETM 2010, Amsterdam, The Netherlands, September 6, 2010. Proceedings, volume 6236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15484-0 (print), 3-642-15485-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15485-0.

Shumaker:2007:VRS

[Shu07] Randall Shumaker, editor. Virtual Reality: Second International Conference, ICVR 2007, Held as part of HCI International 2007, Beijing, China, July 22– 27, 2007. Proceedings, volume 4563 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73334-5 (print), 3-540-73335-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73335-5.

Shumaker:2009:VMR

[Shu09] Randall Shumaker, editor. Virtual and Mixed Reality: Third International Conference, VMR 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009. Proceedings, volume 5622 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02770-9 (print), 3-642-02771-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02771-0.

Shumaker:2011:VMRa

[Shu11a] Randall Shumaker, editor. Virtual and Mixed Reality — New Trends: International Conference, Virtual and Mixed Reality 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part I, volume 6773 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22020-7 (print), 3-642-22021-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22021-0.

Shumaker:2011:VMRb

[Shu11b] Randall Shumaker, editor. Virtual and Mixed Reality — Systems and Applications: International Conference, Virtual and Mixed Reality 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part II, volume 6774 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22023-1 (print), 3-642-22024-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22024-1.

Suzuki:2008:TSC

[SHUH08] Kenji Suzuki, Teruo Higashino, Andreas Ulrich, and Toru Hasegawa, editors. Testing of Software and Communicating Systems: 20th IFIP TC 6/WG 6.1 International Conference, Test-Com 2008 8th International Workshop, FATES 2008 Tokyo, Japan, June 10–13, 2008 Proceedings, volume 5047 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68514-6 (print), 3-540-68524-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68524-1.

Shvartsman:2006:PDS

[Shv06] Mariam Momenzadeh Alexander A. Shvartsman, editor. Principles of Distributed Systems: 10th International Conference, OPODIS 2006, Bordeaux, France, December 12–15, 2006. Proceedings, volume 4305 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49990-3 (print), 3-540-49991-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        49991-6.

Seipel:2009:ADP

[SHW09] Dietmar Seipel, Michael Hanus, and Armin Wolf, editors. Applications of Declarative Programming and Knowledge Management: 17th International Conference, INAP 2007, and 21st Workshop on Logic Programming, WLP 2007, Würzburg, Germany, October 4–6, 2007, Revised Selected Papers, volume 5437 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00674-4 (print), 3-642-00675-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00675-3.

Sha:2006:EUC

[SHX^{+}06] Edwin Sha, Sung-Kook Han, Cheng-Zhong Xu, Moon-Hae Kim, Laurence T. Yang, and Bin Xiao, editors. Embedded and Ubiquitous Computing: International Conference, EUC 2006, Seoul, Korea, August 1–4, 2006. Proceedings, volume 4096 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36679-2 (print), 3-540-36681-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36681-2.

Shyamasundar:1992:FST

[Shy92] Rudrapatna Shyamasundar, editor. Foundations of software technology and theoretical computer science: 12th conference, New Delhi, India, December 18–20, 1992: proceedings, volume 652 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-56287-7 (Berlin), 0-387-56287-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F68 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0652.htm;
 http://www.springerlink.
com/content/978-3-540-
56287-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=652.

Shyamasundar:1993:FST

[Shy93] Rudrapatna Shyamasundar, editor. Foundations of software technology and theoretical computer science: 13th conference, Bombay, India, December 15–17, 1993: proceedings, volume 761 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57529-4 (Berlin), 0-387-57529-4 (New York).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F69 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0761.htm;
 http://www.springerlink.
com/content/978-3-540-
57529-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=761.

“Papers selected for … FST and TCS Conference”–Pref.

Suzuki:2008:FTN

[SHYEF08] Kenji Suzuki, Teruo Higashino, Keiichi Yasumoto, and Khaled El-Fakih, editors. Formal Techniques for Networked and Distributed Systems — FORTE 2008: 28th IFIP WG 6.1 International Conference Tokyo, Japan, June 10–13, 2008 Proceedings, volume 5048 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68854-4 (print), 3-540-68855-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68855-6.

Siefkes:1975:GVJ

[Sie75] Dirk Siefkes, editor. GI-4. [i.e., vierte] Jahrestagung: Berlin, 9.–12. Oktober 1974. (German) [GI-4. Fourth Annual Conference: Berlin, October 9–12, 1974], volume 26 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07141-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.26.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0026.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        07141-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=26.

English or German.

Siekmann:1986:ICA

[Sie86] Jorg H. Siekmann, editor. 8th International Conference on Automated Deduction: Oxford, England, July 27– August 1, 1986: proceedings, volume 230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16780-3 (paperback).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I571 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0230.htm;
 http://www.springerlink.
com/content/978-0-387-
16780-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=230.

Sifakis:1990:AVM

[Sif90] J. Sifakis, editor. Automatic verification methods for finite state systems: International Workshop, Grenoble, France, June 12–14, 1989: proceedings, volume 407 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52148-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .A97 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0407.htm;
 http://www.springerlink.
com/content/978-0-387-
52148-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=407.

Silverman:2001:CLI

[Sil01] Joseph H. Silverman, editor. Cryptography and lattices: International Conference, CaLC 2001, Providence RI, USA, March 29– 30, 2001: Revised Papers, volume 2146 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42488-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .C35 2001; QA267.A1 L43 no.2146. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2146.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42488-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2146.

Shimojo:2005:WCT

[SILS05] Shinji Shimojo, Shingo Ichii, Tok Wang Ling, and Kwan-Ho Song, editors. Web and communication technologies and Internet-related social issues–HSI 2005: 3rd International Conference on Human.Society@Internet, Tokyo, Japan, July 27– 29, 2005: proceedings, volume 3597 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27830-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 I5345 2005. URL http:

//www.springerlink.
com/content/978-3-540-
27830-6;   http://www.
springerlink.com/link.
asp?id=0j4995tm5bh3.

Simon:1992:LLA

[Sim92] Imre Simon, editor.

LATIN ’92: 1st Latin American Symposium on Theoretical Informatics, Sao Paulo, Brazil, April 6–10, 1992: proceedings, volume 583 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55284-7 (Berlin), 0-387-55284-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .L38 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0583.htm;
 http://www.springerlink.
com/content/978-3-540-
55284-0;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=583.

Singh:1994:MST

[Sin94] Munindar P. (Munindar Paul) Singh. Multiagent systems: a theoretical framework for intentions, know-how, and communications, volume 799 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58026-3 (Berlin), 0-387-58026-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xxiii + 168 pp. LCCN QA76.9.D5 S545 1994.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0799.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        58026-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=799.

Satoh:2008:NFA

[SINK08] Ken Satoh, Akihiro Inokuchi, Katashi Nagao, and Takahiro Kawamura, editors. New Frontiers in Artificial Intelligence: JSAI 2007 Conference and Workshops, Miyazaki, Japan, June 18–22, 2007, Revised Selected Papers, volume 4914 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78196-X (print), 3-540-78197-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78197-4.

Sion:2010:FCDa

[Sio10] Radu Sion, editor. Financial Cryptography and Data Security: 14th International Conference, FC 2010, Tenerife, Canary Islands, January 25–28, 2010, Revised Selected Papers, volume 6052 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14576-0 (print), 3-642-14577-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14577-3.

Sipper:1997:EPC

[Sip97]  Moshe Sipper.  Evo-

lution of parallel cellular machines: the cellular programming approach, volume 1194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62613-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 198 pp. LCCN QA76.58.S575 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1194.
htm;   http://www.springerlink.
com/content/978-3-540-
62613-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1194.

Sirjani:2012:CML

[Sir12] Marjan Sirjani, editor. Coordination Models and Languages: 14th International Conference, COORDINATION 2012, Stockholm, Sweden, June 14–15, 2012. Proceedings, volume 7274 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30828-7 (print), 3-642-30829-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
30829-1.

Shi:2006:DWI

[SJ06] Yun Qing Shi and Byeungwoo Jeon, editors. Digital Watermarking: 5th International Workshop, IWDW 2006, Jeju Island, Korea, November 8–10, 2006. Proceedings, volume 4283 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48825-1 (print), 3-540-48827-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48827-9.

Schouten:2008:BIM

[SJDT08] Ben Schouten, Niels Christian Juul, Andrzej Drygajlo, and Massimo Tistarelli, editors. Biometrics and Identity Management: First European Workshop, BIOID 2008, Roskilde, Denmark, May 7–9, 2008. Revised Selected Papers, volume 5372 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89990-1 (print), 3-540-89991-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89991-4.

Setchi:2010:KBIa

[SJHJ10a] Rossitza Setchi, Ivan Jordanov, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 14th International Conference, KES 2010, Cardiff, UK, September 8–10, 2010, Proceedings, Part I, volume 6276 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15386-0 (print), 3-642-15387-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15387-7.

Setchi:2010:KBIb

[SJHJ10b] Rossitza Setchi, Ivan Jordanov, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 14th International Conference, KES 2010, Cardiff, UK, September 8–10, 2010, Proceedings, Part II, volume 6277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15389-5 (print), 3-642-15390-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15390-7.

Setchi:2010:KBIc

[SJHJ10c] Rossitza Setchi, Ivan Jordanov, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 14th International Conference, KES 2010, Cardiff, UK, September 8–10, 2010, Proceedings, Part III, volume 6278 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15392-5 (print), 3-642-15393-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15393-8.

Setchi:2010:KBId

[SJHJ10d] Rossitza Setchi, Ivan Jordanov, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 14th International Conference, KES 2010, Cardiff, UK, September 8–10, 2010, Proceedings, Part IV, volume 6279 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15383-6 (print), 3-642-15384-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15384-6.

Simon:2005:ALT

[SJT05] Hans Ulrich Simon, Sanjay Jain, and Etsuji Tomita, editors. Algorithmic learning theory: 16th international conference, ALT 2005, Singapore, October 8– 11, 2005: proceedings, volume 3734 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29242-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 A48 2005.

URL http://www.springerlink.
com/content/978-3-540-
29242-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3734.

Scholl:2012:EGI

[SJW^{+}12] Hans J. Scholl, Marijn Janssen, Maria A. Wimmer, Carl Erik Moe, and Leif Skiftenes Flak, editors. Electronic Government: 11th IFIP WG 8.5 International Conference, EGOV 2012, Kristiansand, Norway, September 3–6, 2012. Proceedings, volume 7443 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33488-1 (print), 3-642-33489-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33489-4.

Sommer:1997:AFP

[SK97] Gerald Sommer and Jan J. Koenderink, editors. Algebraic frames for the perception-action cycle: International Workshop, AFPAC ’97, Kiel, Germany, September 8– 9, 1997: proceedings, volume 1315 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63517-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ217.5.A39 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1315.
htm;   http://www.springerlink.
com/content/978-3-540-
63517-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1315.

Sommer:2008:RVS

[SK08] Gerald Sommer and Reinhard Klette, editors. Robot Vision: Second International Workshop, RobVis 2008, Auckland, New Zealand, February 18–20, 2008. Proceedings, volume 4931 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78156-0 (print), 3-540-78157-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
78157-8.

SanMartin:2011:PPR

[SK11] César San Martin and Sang-Woon Kim, editors. Progress in Pattern Recognition, Image Analysis, Computer Vision, and Applications: 16th Iberoamerican Congress, CIARP 2011, Pucón, Chile, November 15– 18, 2011. Proceedings, volume 7042 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25084-X (print), 3-642-25085-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25085-9.

Shi:2012:TDHb

[SK12] Yun Q. Shi and Stefan Katzenbeisser, editors. Transactions on Data Hiding and Multimedia Security VIII: Special Issue on Pattern Recognition for IT Security, volume 7228 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31970-X (print), 3-642-31971-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31971-6.

Skarbek:2001:CAI

[Ska01] Wladyslaw Skarbek, editor. Computer analysis of images and patterns: 9th International Conference, CAIP 2001, Warsaw, Poland, September 5–7, 2001: Proceedings, volume 2124 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42513-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637 .I53 2001; QA267.A1 L43 no.2124. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2124.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42513-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2124.

Streitz:1998:CBI

[SKB98] N. A. (Norbert A.) Streitz, Shinichi Konomi, and Heinz-Jürgen Burkhardt, editors. Cooperative buildings: integrating information, organization, and architecture: First International Workshop, CoBuild ’98, Darmstadt, Germany, February 25–26, 1998 proceedings, volume 1370 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64237-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN NA2728.C59 1998. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1370.
htm;   http://www.springerlink.
com/content/978-3-540-
64237-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1370.

Silva:2010:WSN

[SKB10] Jorge Sá Silva, Bhaskar Krishnamachari, and Fernando Boavida, editors. Wireless Sensor Networks: 7th European Conference, EWSN 2010, Coimbra, Portugal, February 17–19, 2010. Proceedings, volume 5970 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11916-6 (print), 3-642-11917-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        11917-0.

Stajano:2007:UCT

[SKCK07] Frank Stajano, Hyoung Joong Kim, Jong-Suk Chae, and Seong-Dong Kim, editors. Ubiquitous Convergence Technology: First International Conference, ICUCT 2006, Jeju Island, Korea, December 5–6, 2006, Revised Selected Papers, volume 4412 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71788-9 (print), 3-540-71789-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71789-8.

Sonka:2004:CVM

[SKK04] Milan Sonka, Ioannis A. Kakadiaris, and Jan Kybic, editors. Computer Vision and Mathematical Methods in Medical and Biomedical Image Analysis: ECCV 2004 Workshops CVAMIA and MMBIA, Prague, Czech Republic, May 15, 2004, Revised Selected Papers, volume 3117 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22675-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3117.htm;
 http://www.springerlink.
com/content/978-3-540-
22675-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3117;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98995.

Shi:2008:DWI

[SKK08] Yun Q. Shi, Hyoung-Joong Kim, and Stefan Katzenbeisser, editors. Digital Watermarking: 6th International Workshop, IWDW 2007 Guangzhou, China, December 3–5, 2007 Proceedings, volume 5041 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92237-7 (print), 3-540-92238-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        92238-4.

Streitz:2007:DCI

[SKM07] Norbert Streitz, Achilles Kameas, and Irene Mavrommati, editors. The Disappearing Computer: Interaction Design, System Infrastructures and Applications for Smart Environments, volume 4500 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72725-6 (print), 3-540-72727-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72727-9.

Salomaa:1994:RTT

[SKMR94] Arto Salomaa, J. Karhumaki, Hermann A. Maurer, and Grzegorz Rozenberg, editors. Results and trends in theoretical computer science: colloquium in honor of Arto Salomaa, Graz, Austria, June 10–11, 1994: proceedings, volume 812 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58131-6 (Berlin), 0-387-58131-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .R46 1994. DM86.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0812.
htm;   http://www.springerlink.
com/content/978-3-540-
58131-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=812.

Schillo:2004:MST

[SKMT04] Michael Schillo, Matthias Klusch, Jörg Müller, and Huaglory Tianfield, editors. Multiagent System Technologies: First German Conference, MATES 2003, Erfurt, Germany, September 22–25, 20003: Proceedings, volume 2831 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20124-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M38 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2831.
htm;   http://www.springerlink.
com/content/978-3-540-
20124-3;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2831;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13850.

Szczepaniak:2005:AWI

[SKN05] Piotr S. Szczepaniak, Janusz Kacprzyk, and Adam Niewiadomski, editors. Advances in Web Intelligence: Third International Atlantic Web Intelligence Conference, AWIC 2005, Łodz, Poland, June 6–9, 2005: proceedings, volume 3528 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26219-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .I5726 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26219-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3528;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b137066.

Skowron:1985:CTF

[Sko85] Andrzej Skowron, editor. Computation theory: fifth symposium, Zaborow, Poland, December 3–8, 1984: proceedings, volume 208 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-16066-3 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .C58671 1984. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0208.htm;
 http://www.springerlink.
com/content/978-0-387-
16066-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=208.

Papers from the 5th Symposium on Computation Theory.

Sojka:2000:TSD

[SKP00] Petr Sojka, Ivan Kopecek, and Karel Pala, editors. Text, Speech and Dialogue: Third International Workshop, TSD 2000, Brno, Czech Republic, September 13– 16, 2000. Proceedings, volume 1902 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41042-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 T47 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1902.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41042-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1902.

Sojka:2002:TSD

[SKP02] Petr Sojka, Ivan Kopecek, and Karel Pala, editors. Text, Speech and Dialogue: 5th International Conference, TSD 2002, Brno, Czech Republic, September 9–12, 2002. Proceedings, volume 2448 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44129-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 T47 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2448.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44129-8;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2448.

Sojka:2004:TSD

[SKP04] Petr Sojka, Ivan Kopecek, and Karel Pala, editors. Text, Speech and Dialogue: 7th International Conference, TSD 2004, Brno, Czech Republic, September 8–11, 2004. Proceedings, volume 3206 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23049-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3206.htm;
http://www.springerlink.
com/content/978-3-540-
23049-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3206;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100511.

Sojka:2006:TSD

[SKP06] Petr Sojka, Ivan Kopeček, and Karel Pala, editors. Text, Speech and Dialogue: 9th International Conference, TSD 2006, Brno, Czech Republic, September 11–15, 2006.

Proceedings, volume 4188 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39090-1 (print), 3-540-39091-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
39091-6.

Schilder:2007:AER

[SKP07] Frank Schilder, Graham Katz, and James Pustejovsky, editors. Annotating, Extracting and Reasoning about Time and Events: International Seminar, Dagstuhl Castle, Germany, April 10–15, 2005. Revised Papers, volume 4795 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75988-3 (print), 3-540-75989-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75989-8.

Schmid:2010:AAI

[SKP10] Ute Schmid, Emanuel Kitzelmann, and Rinus Plasmeijer, editors. Approaches and Applications of Inductive Programming: Third International Workshop, AAIP 2009, Edinburgh, UK, September 4, 2009. Revised Papers, volume 5812 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11930-1 (print), 3-642-11931-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11931-6.

Shi:2012:DFW

[SKPG12] Yun Qing Shi, Hyoung-Joong Kim, and Fernando Perez-Gonzalez, editors. Digital Forensics and Watermarking: 10th International Workshop, IWDW 2011, Atlantic City, NY, October 23–26, 2011, Revised Selected Papers, volume 7128 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32204-2 (print), 3-642-32205-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32205-1.

Szczuka:2010:RSC

[SKR^{+}10] Marcin Szczuka, Marzena Kryszkiewicz, Sheela Ramanna, Richard Jensen, and Qinghua Hu, editors. Rough Sets and Current Trends in Computing: 7th International Conference, RSCTC 2010, Warsaw, Poland, June 28–30,2010. Proceedings, volume 6086 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13528-5 (print), 3-642-13529-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13529-3.

Sansone:2011:MCS

[SKR11] Carlo Sansone, Josef Kittler, and Fabio Roli, editors. Multiple Classifier Systems: 10th International Workshop, MCS 2011, Naples, Italy, June 15–17, 2011. Proceedings, volume 6713 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21556-4 (print), 3-642-21557-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21557-5.

Serugendo:2004:ESO

[SKRZ04] Giovanna Di Marzo Serugendo, Anthony Karageorgos, Omer F. Rana, and Franco Zambonelli, editors. Engineering Self-Organising Systems: Nature-Inspired Approaches to Software Engineering, volume 2977 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21201-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E54 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2977.htm;
 http://www.springerlink.
com/content/978-3-540-
21201-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2977;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95863.

Sahni:2002:HPC

[SKS02] Sartaj Sahni, V. K. Prasanna Kumar, and Uday Shukla, editors. High performance computing, HiPC 2002: 9th International Conference, Bangalore, India, December 18– 21, 2002: Proceedings, volume 2552 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00303-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I5847 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2552.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00303-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2552.

Also available via the

World Wide Web.

Sakakibara:2006:GIA

[SKS^{+}06] Yasubumi Sakakibara, Satoshi Kobayashi, Kengo Sato, Tetsuro Nishino, and Etsuji Tomita, editors. Grammatical Inference: Algorithms and Applications: 8th International Colloquium, ICGI 2006, Tokyo, Japan, September 20–22, 2006.

Proceedings, volume 4201 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45264-8 (print), 3-540-45265-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45265-2.

Sendhoff:2009:CBL

[SKS^{+}09] Bernhard Sendhoff, Edgar Körner, Olaf Sporns, Helge Ritter, and Kenji Doya, editors. Creating Brain-Like Intelligence: From Basic Principles to Complex Intelligent Systems, volume 5436 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00615-9 (print), 3-642-00616-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00616-6.

Shibuya:2012:PRB

[SKSA12] Tetsuo Shibuya, Hisashi Kashima, Jun Sese, and Shandar Ahmad, editors. Pattern Recognition in Bioinformatics:

7th IAPR International Conference, PRIB 2012, Tokyo, Japan, November 8–10, 2012. Proceedings, volume 7632 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34122-5 (print), 3-642-34123-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34123-6.

Song:2006:ISA

[SKY06] JooSeok Song, Taekyoung Kwon, and Moti Yung, editors. Information security applications: 6th international workshop, WISA 2005, Jeju Island, Korea, August 22–24, 2005: revised selected papers, volume 3786 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31012-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 W577 2006.

URL http://www.springerlink.
com/content/978-3-540-
31012-9;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3786.

Sivasubramaniam:1999:NBP

[SL99a] Anand Sivasubramaniam and Mario Lauria, editors. Network-based parallel computing: communication, architecture, and applications: Third International Workshop, CANPC’99, Orlando, Florida, USA, January 9th, 1999: Proceedings, volume 1602 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65915-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .C37 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1602.htm;
http://www.springerlink.
com/content/978-3-540-
65915-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1602.

Solina:1999:CAI

[SL99b] Franc Solina and Ales Leonardis, editors. Computer analysis of images and patterns: 8th International Conference, CAIP’99, Ljubliana, Slovenia, September 1–3, 1999: Proceedings, volume 1689 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66366-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1689; TA1637 .C66 1999; TA1632 .I547 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1689.htm;
 http://www.springerlink.
com/content/978-3-540-
66366-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1689.

Steffen:2003:VMC

[SL03] Bernhard Steffen and Giorgio Levi, editors. Verification, Model Checking, and Abstract Interpretation: 5th International Conference, VM-CAI 2004 Venice, Italy, January 11–13, 2004 Proceedings, volume 2937 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20803-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 V53 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2937.
htm;   http://www.springerlink.

com/content/978-3-540-
20803-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2937;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94790.

Sanfeliu:2005:PPR

[SL05] Alberto Sanfeliu and Manuel Lazo Cortés, editors. Progress in pattern recognition, image analysis and applications: 10th Iberoamerican Congress on Pattern Recognition, CIARP 2005, Havana, Cuba, November 15–18, 2005: proceedings, volume 3773 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29850-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 I198 2005.

URL http://www.springerlink.
com/content/978-3-540-
29850-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3773.

Salah:2011:HBU

[SL11] Albert Ali Salah and Bruno Lepri, editors. Human Behavior Understanding: Second International Workshop, HBU 2011, Amsterdam, The Netherlands, November 16, 2011. Proceedings, volume 7065 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25445-4 (print), 3-642-25446-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25446-8.

Salampasis:2012:MIR

[SL12] Michail Salampasis and Birger Larsen, editors. Multidisciplinary Information Retrieval: 5th International Retrieval Facility Conference, IRFC 2012, Vienna, Austria, July 2–3, 2012 Proceedings, volume 7356 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31273-X (print), 3-642-31274-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31274-8.

Shen:2005:CSC

[SLB05] Weiming Shen, Zongkai Lin, and Jean-Paul A. Barthes, editors. Computer supported cooperative work in design I: 8th international conference, CSCWD 2004, Xiamen, China, May 26– 28, 2004: revised selected papers, volume 3168 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29400-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA174 .I567 2004. URL http:

//www.springerlink.
com/content/978-3-540-
29400-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3168.

Shum:2001:AMI

[SLC01] Heung-Yeung Shum, Mark Liao, and Shih-Fu Chang, editors. Advances in multimedia information processing-PCM 2001: second IEEE Pacific Rim Conference on Multimedia, Beijing, China, October 24–26, 2001: Proceedings, volume 2195 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42680-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I333 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2195.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42680-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2195.

Sun:2011:BRC

[SLCT11] Zhenan Sun, Jianhuang Lai, Xilin Chen, and Tieniu Tan, editors. Biometric Recognition: 6th Chinese Conference, CCBR 2011, Beijing, China, December 3–4, 2011. Proceedings, volume 7098 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25448-9 (print), 3-642-25449-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25449-9.

Sledge:1992:SEE

[Sle92] C. Sledge, editor. Software engineering education: SEI Conference 1992, San Diego, California, USA, October 5–7, 1992: proceedings, volume 640 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55963-9 (Berlin), 0-387-55963-9 (New York Heidelberg Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S64625 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0640.
htm;   http://www.springerlink.
com/content/978-3-540-
55963-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=640.

Proceedings of the Sixth SEI Conference on Software Engineering Education.

Sriram:1991:CAC

[SLF91] D. Sriram, Robert D.

Logcher, and S. Fukuda, editors. Computer-aided cooperative product development: MIT-JSME workshop, MIT, Cambridge, USA, November 20–21, 1989: proceedings, volume 492 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54008-3 (New York), 3-540-54008-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .C634 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0492.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        54008-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=492.

Selection of papers presented at the MIT-JSME Workshop on Cooperative Product Development, held Nov. 20– 21, 1989, at the Massachusetts Institute of Technology.

Sebe:2004:CVH

[SLH04] Nicu Sebe, Michael S. Lew, and Thomas S. Huang, editors. Computer Vision in Human-Computer Interaction: ECCV 2004 Workshop on HCI, Prague, Czech Republic, May 16, 2004: Proceedings, volume 3058 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22012-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 E33 2004.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3058.

htm;   http://www.springerlink.
com/content/978-3-540-
22012-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3058;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97917.

Sebe:2005:CVH

[SLH05] Nicu Sebe, Michael S.

Lew, and Thomas S. Huang, editors. Computer vision in humancomputer interaction: ICCV 2005 Workshop on HCI, Beijing, China, October 21, 2005: proceedings, volume 3766 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29620-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I617 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29620-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3766.

Spaniol:2009:AWB

[SLKL09] Marc Spaniol, Qing Li, Ralf Klamma, and Rynson W. H. Lau, editors. Advances in Web Based Learning — ICWL 2009:

8th International Conference, Aachen, Germany, August 19–21, 2009. Proceedings, volume 5686 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03425-X (print), 3-642-03426-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03426-8.

Sloman:2001:PDS

[SLL01] Morris Sloman, Jorge Lobo, and Emil C. Lupu, editors. Policies for distributed systems and networks: International Workshop, POL-ICY 2001, Bristol, UK, January 29–31, 2001: Proceedings, volume 1995 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41610-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 P63 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1995.htm;
 http://www.springerlink.
com/content/978-3-540-
41610-4;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1995.

Shen:2006:AWN

[SLL^{+}06] Heng Tao Shen, Jinbao Li, Minglu Li, Jun Ni, and Wei Wang, editors. Advanced Web and network technologies, and applications: APWeb 2006 international workshops: XRA, IWSN, MEGA, and ICSE, Harbin, China, January 16–18, 2006: proceedings, volume 3842 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31158-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .A7852 2006.

URL http://www.springerlink.
com/content/978-3-540-
31158-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3842.

Shen:2007:CSC

[SLL^{+}07] Weiming Shen, Junzhou Luo, Zongkai Lin, Jean-Paul A. Barthès, and Qi Hao, editors. Computer Supported Cooperative Work in Design III: 10th International Conference, CSCWD 2006, Nanjing, China, May 3– 5, 2006, Revised Selected Papers, volume 4402 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72862-7 (print), 3-540-72863-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72863-4.

Song:2003:CMI

[SLLS03] Il-Yeol Song, Stephen W. Liddle, Tok Wang Ling, and Peter Scheuermann, editors. Conceptual Modeling—ER 2003: 22nd International Conference on Conceptual Modeling, Chicago, IL, USA, October 13–16, 2003: Proceedings, volume 2813 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20299-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2813.htm;
 http://www.springerlink.
com/content/978-3-540-
20299-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2813;

 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13244.

Susilo:2007:PSF

[SLM07] Willy Susilo, Joseph K.

Liu, and Yi Mu, editors. Provable Security: First International Conference, ProvSec 2007, Wollongong, Australia, November 1–2, 2007. Proceedings, volume 4784 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75669-8 (print), 3-540-75670-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75670-5.

Sloot:1999:HPC

[Slo99] Peter Sloot, editor. Highperformance computing and networking: 7th International Conference, HPCN Europe, 1999, Amsterdam, The Netherlands, April 1999: Proceedings, volume 1593 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65821-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H52 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1593.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65821-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1593.

Strang:2005:LCA

[SLP05] Thomas Strang and Claudia Linnhoff-Popien, editors. Location- and context-awareness: first international workshop, LOCA 2005, Oberpfaffenhofen, Germany, May 12–13, 2005, proceedings, volume 3479 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25896-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25896-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3479;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136418.

Spaniol:1996:TDS

[SLPM96] Otto Spaniol, Claudia Linnhoff-Popien, and Bernd E. Meyer, editors. Trends in distributed systems: CORBA and beyond: International Workshop TreDS ’96, Aachen, Germany, October 1–2, 1996: proceedings, volume 1161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61842-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I63 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1161.htm;
http://www.springerlink.
com/content/978-3-540-
61842-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1161.

Snoeyink:2012:FAA

[SLSW12] Jack Snoeyink, Pinyan Lu, Kaile Su, and Lusheng Wang, editors. Frontiers in Algorithmics and Algorithmic Aspects in Information and Management: Joint International Conference, FAW-AAIM 2012, Beijing, China, May 14–16, 2012. Proceedings, volume 7285 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29699-8 (print), 3-642-29700-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29700-7.

Sebe:2007:MCA

[SLZH07] Nicu Sebe, Yuncai Liu, Yueting Zhuang, and Thomas S. Huang, editors. Multimedia Content Analysis and Mining: International Workshop, MCAM 2007, Weihai, China, June 30– July 1, 2007. Proceedings, volume 4577 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73416-3 (print), 3-540-73417-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73417-8.

Schwefel:1991:PPS

[SM91] Hans-Paul Schwefel and Reinhard Manner, editors. Parallel problem solving from nature:

1st workshop, PPSN I, Dortmund, FRG, October 1–3, 1990: proceedings, volume 496 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54148-9 (Berlin), 0-387-54148-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P3777 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0496.htm;
 http://www.springerlink.
com/content/978-3-540-
54148-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=496.

Stroulia:2001:AAI

[SM01] Eleni Stroulia and Stan Matwin, editors. Advances in artificial intelligence: 14th Biennial Conference of the Canadian Society for Computational Studies of Intelligence, AI 2001, Ottawa, Canada, June 7–9, 2001: Proceedings, volume 2056 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42144-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2056; Q334 .C36 2001. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2056.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        42144-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2056.

Schwenker:2006:ANN

[SM06]  Friedhelm  Schwenker

and Simone Marinai, editors. Artificial Neural Networks in Pattern Recognition: Second IAPR Workshop, ANNPR 2006, Ulm, Germany, August 31– September 2, 2006. Proceedings, volume 4087 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37951-7 (print), 3-540-37952-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37952-2.

Svensson:2009:ICS

[SM09] Lars Svensson and José Monteiro, editors. Integrated Circuit and System Design. Power and Timing Modeling, Optimization and Simulation: 18th International Workshop, PAT-MOS 2008, Lisbon, Portugal, September 10–12, 2008. Revised Selected Papers, volume 5349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-95947-5 (print), 3-540-95948-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
95948-9.

Sakakibara:2011:DCM

[SM11a] Yasubumi Sakakibara and Yongli Mi, editors. DNA Computing and Molecular Programming: 16th International Conference, DNA 16, Hong Kong, China, June 14–17, 2010, Revised Selected Papers, volume 6518 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18304-2 (print), 3-642-18305-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        18305-8.

Simao:2011:FMF

[SM11b] Adenilso Simao and Carroll Morgan, editors. Formal Methods, Foundations and Applications: 14th Brazilian Symposium, SBMF 2011, São Paulo, Brazil, September 26–30, 2011, Revised Selected Papers, volume 7021 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25031-9 (print), 3-642-25032-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25032-3.

Small:1997:ADB

[Sma97] Carol Small, editor. Advances in databases: 15th British National Conference on Databases, BNCOD 15, London, United Kingdom, July 7–9, 1997: proceedings, volume 1271 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63263-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25A279 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1271.
htm;   http://www.springerlink.
com/content/978-3-540-
63263-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1271.

Spaccapietra:2003:JDSa

[SMA03] Stefano Spaccapietra, Sal March, and Karl Aberer, editors. Journal on Data Semantics. I, volume 2800 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20407-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 232 pp. LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2800.htm;
http://www.springerlink.
com/content/978-3-540-
20407-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2800;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b14200.

Smart:2005:CCI

[Sma05] Nigel P. Smart, editor. Cryptography and Coding: 10th IMA international Conference, Cirencester, UK, December 19–21, 2005. Proceedings, volume 3796 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30276-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30276-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3796.

Smart:2008:ACE

[Sma08] Nigel Smart, editor. Advances in Cryptology

— EUROCRYPT 2008: 27th Annual International Conference on the Theory and Applications of Cryptographic Techniques, Istanbul, Turkey, April 13–17, 2008. Proceedings, volume 4965 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. ISBN 3-540-78967-7. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78967-3.

Schaeffer:2003:CGT

[SMB03] Jonathan Schaeffer, Martin Müller, and Yngvi Björnsson, editors. Computers and Games: Third International Conference, CG 2002, Edmonton, Canada, July 25–27, 2002: Revised Papers, volume 2883 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20545-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 C4166 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2883.htm;
http://www.springerlink.
com/content/978-3-540-
20545-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2883;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94154.

Sauveron:2007:IST

[SMBQ07] Damien Sauveron, Konstantinos Markantonakis, Angelos Bilas, and Jean-Jacques Quisquater, editors. Information Security Theory and Practices. Smart Cards, Mobile and Ubiquitous Computing Systems: First IFIP TC6 /WG 8.8 / WG 11.2 International Workshop, WISTP 2007, Heraklion, Crete, Greece, May 9–11, 2007. Proceedings, volume 4462 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72353-6 (print), 3-540-72354-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72354-7.

Stajano:2007:SPA

[SMCM07] Frank Stajano, Catherine Meadows, Srdjan Capkun, and Tyler Moore, editors. Security and Privacy in Ad-hoc and Sensor Networks: 4th European Workshop, ESAS 2007, Cambridge, UK, July 2–3, 2007. Proceedings, volume 4572 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73274-8 (print), 3-540-73275-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73275-4.

Sahinalp:2004:CPM

[SMD04] Suleyman Cenk Sahinalp, S. Muthukrishnan, and Ugur Dogrusoz, editors. Combinatorial Pattern Matching: 15th annual symposium, CPM 2004, Istanbul, Turkey, July 5–7, 2004: Proceedings, volume 3109 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22341-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3109.
htm;   http://www.springerlink.
com/content/978-3-540-
22341-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3109;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98377.

Sirisuk:2010:RCA

[SMEGA10] Phaophak Sirisuk, Fearghal Morgan, Tarek El-Ghazawi, and Hideharu Amano, editors. Reconfigurable Computing: Architectures, Tools and Applications: 6th International Symposium, ARC 2010, Bangkok, Thailand, March 17–19, 2010. Proceedings, volume 5992 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12132-2 (print), 3-642-12133-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12133-3.

Song:2011:QII

[SMF^{+}11] Dawei Song, Massimo Melucci, Ingo Frommholz, Peng Zhang, Lei Wang, and Sachi Arafat, editors. Quantum Interaction: 5th International Symposium, QI 2011, Aberdeen, UK, June 26–29, 2011, Revised Selected Papers, volume 7052 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24970-1 (print), 3-642-24971-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24971-6.

Schoeffmann:2012:AMM

[SMH^{+}12] Klaus Schoeffmann, Bernard Merialdo, Alexander G. Hauptmann, Chong-Wah Ngo, Yiannis Andreopoulos, and Christian Breiteneder, editors. Advances in Multimedia Modeling: 18th International Conference, MMM 2012, Klagenfurt, Austria, January 4–6, 2012. Proceedings, volume 7131 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27354-8 (print), 3-642-27355-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
27355-1.

Smith:1974:MER

[Smi74] Brian T. Smith. Matrix eigensystem routines. EISPACK guide, volume 6 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1974. CODEN LNCSD9. ISBN 0-387-06710-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 387 pp. LCCN QA267.A1 L43 no.6.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0006.
htm;   http://www.springerlink.

                                        com/content/978-0-387-
                                        06710-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=6.

Smith:1998:AIS

[Smi98] Ian Smith. Artificial intelligence in structural engineering: information technology for design, collaboration, maintenance, and monitoring, volume 1454 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64806-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 496 pp. LCCN TA636 .A77 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1454.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64806-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1455.

Smirnov:2001:QFI

[Smi01] Mikhail I. Smirnov, editor. Quality of future Internet services: Second COST 263 International Workshop, QofIS 2001, Coimbra, Portugal, September 24–26, 2001:

Proceedings, volume 2156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42602-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2156; TK5105.875.I57 C67 2001; TK5105.875.I57 Q36 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2156.htm;
 http://www.springerlink.
com/content/978-3-540-
42602-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2156.

Smirnov:2003:QFI

[Smi03] Michael Smirnov, editor. Quality of Future Internet Services: COST Action 263 Final Report, volume 2856 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20193-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 Q34 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2856.
htm;   http://www.springerlink.
com/content/978-3-540-
20193-9;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2856;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13823.

Smirnov:2005:ACF

[Smi05] Michael Smirnov, editor. Autonomic Communication: First International IFIP Workshop, WAC 2004, Berlin, Germany, October 18–19, 2004, Revised Selected Papers, volume 3457 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27417-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A97 W33 2004.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        27417-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3457.

Smith:2006:ICE

[Smi06] Ian F. C. Smith, editor. Intelligent Computing in Engineering and Architecture: 13th EG-ICE Workshop 2006, Ascona, Switzerland, June 25–30, 2006, Revised Selected Papers, volume 4200 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46246-5 (print), 3-540-46247-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46247-7.

Smith:2012:ITS

[Smi12] Adam Smith, editor. Information Theoretic Security: 6th International Conference, ICITS 2012, Montreal, QC, Canada, August 15–17, 2012. Proceedings, volume 7412 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32283-2 (print), 3-642-32284-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32284-6.

Singh:2001:APR

[SMK01] Sameer Singh, Nabeel A.

Murshed, and Walter Kropatsch, editors. Advances in pattern recognition: ICAPR 2001: Second International Conference, Rio de Janeiro, Brazil, March 11–14, 2001: Proceedings, volume 2013 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41767-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 A34 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2013.htm;
 http://www.springerlink.
com/content/978-3-540-
41767-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2013.

Spaccapietra:2002:CMI

[SMK02] S. Spaccapietra, Salvatore T. March, and Y. Kambayashi, editors. Conceptual Modeling— ER 2002: 21st International Conference on Conceptual Modeling Tampere, Finland, October 7–11, 2002. Proceedings, volume 2503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44277-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26 I567 2002. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t2503.htm;
http://www.springerlink.
com/content/978-3-540-
44277-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2503.

Also available via the World Wide Web.

Spirakis:2006:INE

[SMK06] Paul Spirakis, Marios Mavronicolas, and Spyros Kontogiannis, editors. Internet and Network Economics: Second International Workshop, WINE 2006, Patras, Greece, December 15–17, 2006. Proceedings, volume 4286 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-68138-8 (print), 3-540-68141-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68141-0.

Sorevik:2001:APC

[SMMG01] Tor Sørevik, Fredrik Manne, Randi Moe, and Assefaw Hadish Gebremedhin, editors. Applied parallel computing: new paradigms for HPC in industry and academia: 5th International Workshop, PARA 2000, Bergen, Norway, June 18–20, 2000: Proceedings, volume 1947 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41729-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P353 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1947.htm;
 http://www.springerlink.
com/content/978-3-540-
41729-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1947.

Smolka:1997:PPC

[Smo97] Gert Smolka, editor. Principles and practice of constraint programming — CP 97: Third International Conference, CP 97, Linz, Austria, October 29– November 1, 1997: proceedings, volume 1330 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63753-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612.C7 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1330.
htm;   http://www.springerlink.
com/content/978-3-540-
63753-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1330.

Smolka:2000:PLS

[Smo00] Gert Smolka, editor. Programming languages and systems: 9th European Symposium on Programming, ESOP 2000, held as part of the Joint European Conference on Theory and Practice of Software, ETAPS 2000, Berlin, Germany, March 25–April 2, 2000: Proceedings, volume 1782 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67262-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1782. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1782.htm;
 http://www.springerlink.
com/content/978-3-540-
67262-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1782.

Sgallari:2007:SSV

[SMP07] Fiorella Sgallari, Almerico Murli, and Nikos Paragios, editors. Scale Space and Variational Methods in Computer Vision: First International Conference, SSVM 2007, Ischia, Italy, May 30 — June 2, 2007. Proceedings, volume 4485 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72822-8 (print), 3-540-72823-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72823-8.

Sipper:1998:ESB

[SMPU98] Moshe Sipper, Daniel Mange, and Andres Pérez-Uribe, editors. Evolvable systems: from biology to hardware: Second International Conference, ICES 98, Lausanne, Switzerland, September 23–25, 1998: Proceedings, volume 1478 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64954-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1478. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1478.htm;
http://www.springerlink.
com/content/978-3-540-
64954-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1478.

Susilo:2012:ISP

[SMS12] Willy Susilo, Yi Mu, and Jennifer Seberry, editors. Information Security and Privacy: 17th Australasian Conference, ACISP 2012, Wollongong, NSW, Australia, July 9–11, 2012. Proceedings, volume 7372 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31447-3 (print), 3-642-31448-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31448-3.

Safavi-Naini:2008:ITS

[SN08] Reihaneh Safavi-Naini, editor. Information Theoretic Security: Third International Conference, ICITS 2008, Calgary, Canada, August 10–13, 2008. Proceedings, volume 5155 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85092-9 (print), 3-540-85093-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85093-9.

Sadre:2012:DNS

[SNČ^{+}12a] Ramin Sadre, Jiří Novotný, Pavel Čeleda, Martin Waldburger, and Burkhard Stiller, editors. Dependable Networks and Services: 6th IFIP WG 6.6 International Conference on Autonomous Infrastructure, Management, and Security, AIMS 2012, Luxembourg, Luxembourg, June 4–8, 2012. Proceedings, volume 7279 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30632-2 (print), 3-642-30633-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30633-4.

Safavi-Naini:2012:ACC

[SNC12b] Reihaneh Safavi-Naini and Ran Canetti, editors. Advances in Cryptology — CRYPTO 2012: 32nd Annual Cryptology Conference, Santa Barbara, CA, USA, August 19–23, 2012. Proceedings, volume 7417 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32008-2 (print), 3-642-32009-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32009-5.

Satoh:2008:AMM

[SNE08] Shin’ichi Satoh, Frank Nack, and Minoru Etoh, editors. Advances in Multimedia Modeling: 14th International Multimedia Modeling Conference, MMM 2008, Kyoto, Japan, January 9– 11, 2008. Proceedings, volume 4903 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77407-6 (print), 3-540-77409-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77409-9.

Schreck:2011:ADG

[SNRG11] Pascal Schreck, Julien Narboux, and Jürgen Richter-Gebert, editors. Automated Deduction in Geometry: 8th International Workshop, ADG 2010, Munich, Germany, July 22–24, 2010, Revised Selected Papers, volume 6877 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25069-6 (print), 3-642-25070-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25070-5.

Safavi-Naini:2003:ISP

[SNS03] Reihanah Safavi-Naini and Jennifer Seberry, editors. Information Security and Privacy: 8th Australasian Conference, ACISP 2003, Wollongong, Australia, July 9–11, 2003: Proceedings, volume 2727 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40515-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 A278 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2727.htm;
http://www.springerlink.
com/content/978-3-540-
40515-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2727.

Sundaram:2006:IVR

[SNSR06] Hari Sundaram, Milind Naphade, John R. Smith, and Yong Rui, editors. Image and Video Retrieval: 5th International Conference, CIVR 2006, Tempe, AZ, USA, July 13–15, 2006. Proceedings, volume 4071 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36018-2 (print), 3-540-36019-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36019-3.

Safavi-Naini:2006:DRM

[SNY06] Reihaneh Safavi-Naini and Moti Yung, editors. Digital Rights Management. Technologies, Issues, Challenges and Systems: First International Conference, DRMTICS 2005, Sydney, Australia, October 31 — November 2, 2005. Revised Selected Papers, volume 3919 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35998-2 (print), 3-540-35999-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35999-9.

Schurr:2008:AGT

[SNZ08] Andy Schürr, Manfred Nagl, and Albert Zündorf, editors. Applications of Graph Transformations with Industrial Relevance: Third International Symposium, AGTIVE 2007, Kassel, Germany, October 10–12, 2007, Revised Selected and Invited Papers, volume 5088 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89019-X (print), 3-540-89020-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89020-1.

Steffen:1999:CSD

[SO99] Bernhard Steffen and E.- R. Olderog, editors. Correct System Design: Recent Insights and Advances, volume 1710 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66624-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 415 pp. LCCN QA76.9.S88 C678 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1710.
htm;   http://www.springerlink.
com/content/978-3-540-
66624-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1710.

Saglietti:2007:CSR

[SO07] Francesca Saglietti and Norbert Oster, editors. Computer Safety, Reliability, and Security: 26th International Conference, SAFECOMP 2007, Nuremberg, Germany, September 18–21, 2007. Proceedings, volume 4680 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75100-9 (print), 3-540-75101-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75101-4.

Senivongse:2009:DAI

[SO09] Twittie Senivongse and Rui Oliveira, editors. Distributed Applications and Interoperable Systems: 9th IFIP WG 6.1 International Conference, DAIS 2009, Lisbon, Portugal, June 9– 11, 2009. Proceedings, volume 5523 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02163-8 (print), 3-642-02164-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02164-0.

Sommerville:1996:SCM

[Som96] Ian Sommerville, editor. Software configuration management: ICSE 96 SCM-6 Workshop, Berlin, Germany, March 1996: Selected Papers, volume 1167 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61964-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C69I25 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1167.
htm;   http://www.springerlink.
com/content/978-3-540-
61964-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1167.

Sommaruga:2009:FTI

[Som09] Giovanni Sommaruga, editor. Formal Theories of Information: From Shannon to Semantic Information Theory and General Concepts of Information, volume 5363 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00658-2 (print), 3-642-00659-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00659-3.

Song:2000:ISC

[Son00] JooSeok Song, editor. Information security and cryptology — ICISC’99: Second International Conference, Seoul, Korea, December 9–10, 1999: Proceedings, volume 1787 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67380-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1787. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1787.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67380-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1787.

Solis-Oba:2012:AOA

[SOP12] Roberto Solis-Oba and Giuseppe Persiano, editors. Approximation and Online Algorithms: 9th International Workshop, WAOA 2011, Saarbrücken, Germany, September 8–9, 2011, Revised Selected Papers, volume 7164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29115-5 (print), 3-642-29116-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29116-6.

Shadbolt:1996:AKA

[SOS96] Nigel Shadbolt, Kieron O’Hara, and Guus Schreiber, editors. Advances in knowledge acquisition: 9th European Knowledge Acquisition Workshop, EKAW ’96, Nottingham, United Kingdom, May 14–17, 1996: proceedings, volume 1076 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61273-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.E95 E92 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1076.htm;
 http://www.springerlink.
com/content/978-3-540-
61273-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1076.

Seberry:1990:ACA

[SP90] Jennifer Seberry and Josef Pieprzyk, editors. Advances in cryptology — AUSCRYPT ’90: international conference on cryptology, Sydney, Australia, January 8–11, 1990: proceedings, volume 453 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-53000-2 (Berlin), 0-387-53000-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 A87 1990. DM63.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0453.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        53000-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=453.

Sponsored by the International Association for Cryptologic Research.

Sommerville:1993:SEE

[SP93] Ian Sommerville and Manfred Paul, editors. Software engineering — ESEC ’93: 4th European Software Engineering Conference, Garmisch-Partenkirchen, Germany, September 13–17, 1993: proceedings, volume 717 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57209-0 (Berlin), 0-387-57209-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .E97 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0717.htm;
 http://www.springerlink.
com/content/978-3-540-
57209-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=717.

Stephanidis:2007:UAA

[SP07] Constantine Stephanidis and Michael Pieper, editors. Universal Access in Ambient Intelligence Environments: 9th ERCIM Workshop on User Interfaces for All, Königswinter, Germany, September 27– 28, 2006. Revised Papers, volume 4397 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71024-8 (print), 3-540-71025-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71025-7.

Sekar:2008:ISS

[SP08] R. Sekar and Arun K. Pujari, editors. Information Systems Security: 4th International Conference, ICISS 2008, Hyderabad, India, December 16–20, 2008. Proceedings, volume 5352 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89861-1 (print), 3-540-89862-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89862-7.

Sadre:2009:SNS

[SP09] Ramin Sadre and Aiko Pras, editors. Scalability of Networks and Services: Third International Conference on Autonomous Infrastructure, Management and Security, AIMS 2009 Enschede, The Netherlands, June 30–July 2, 2009. Proceedings, volume 5637 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02626-5 (print), 3-642-02627-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02627-0.

Stoelinga:2012:FMI

[SP12] Mariëlle Stoelinga and Ralf Pinger, editors. Formal Methods for Industrial Critical Systems: 17th International Workshop, FMICS 2012, Paris, France, August 27–28, 2012. Proceedings, volume 7437 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32468-1 (print), 3-642-32469-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32469-7.

Spaccapietra:2003:JDSb

[Spa03] Stefano Spaccapietra.

Journal on Data Semantics. IV, volume 3730 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-31001-0. ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN QA76.76.I57 J68. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        31001-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3730.

Spaccapietra:2006:JDSc

[Spa06] Stefano Spaccapietra, editor. Journal on Data Semantics VII, volume 4244 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46329-1 (print), 3-540-46330-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46330-6.

Spaccapietra:2008:JDSa

[Spa08] Stefano Spaccapietra, editor. Journal on Data Semantics X, volume 4900 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77687-7 (print), 3-540-77688-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-540-
77688-8.

Spaccapietra:2009:JDSa

[Spa09] Stefano Spaccapietra, editor. Journal on Data Semantics XII, volume 5480 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00684-1 (print), 3-642-00685-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00685-2.

Spaccapietra:2011:JDS

[Spa11] Stefano Spaccapietra, editor. Journal on Data Semantics XV, volume 6720 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22629-9 (print), 3-642-22630-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22630-4.

Soudris:2000:ICD

[SPB00] Dimitrios Soudris, P. Pirsch, and Erich Barke, editors. Integrated Circuit Design. Power and Timing Modeling, Optimization and Simulation: 10th International Workshop, PATMOS 2000, Göttingen, Germany, September 13–15, 2000. Proceedings, volume 1918 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41068-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1918. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1918.htm;
 http://www.springerlink.
com/content/978-3-540-
41068-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1918.

Sadayappan:2008:HPC

[SPBP08] Ponnuswamy Sadayappan, Manish Parashar, Ramamurthy Badrinath, and Viktor K. Prasanna, editors. High Performance Computing — HiPC 2008: 15th International Conference, Bangalore, India, December 17–20, 2008. Proceedings, volume 5374 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89893-X (print), 3-540-89894-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89894-8.

Song:2008:ACM

[SPC^{+}08] Il-Yeol Song, Mario Piattini, Yi-Ping Phoebe Chen, Sven Hartmann, Fabio Grandi, Juan Trujillo, Andreas L. Opdahl, Fernando Ferri, Patrizia Grifoni, Maria Chiara Caschera, Colette Rolland, Carson Woo, Camille Salinesi, Esteban Zimányi, Christophe Claramunt, Flavius Frasincar, Geert-Jan Houben, and Philippe Thiran, editors. Advances in Conceptual Modeling — Challenges and Opportunities: ER 2008 Workshops CMLSA, ECDM, FP-UML, M2AS, RIGiM, SeCoGIS, WISM, Barcelona Spain, October 20–23, 2008. Proceedings, volume 5232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87990-0 (print), 3-540-87991-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87991-6.

Sandoval:2007:CAI

[SPCG07] Francisco Sandoval, Alberto Prieto, Joan Cabestany, and Manuel Graña, editors. Computational and Ambient Intelligence: 9th International Work-Conference on Artificial Neural Networks, IWANN 2007, San Sebastián, Spain, June 20– 22, 2007. Proceedings, volume 4507 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73006-0 (print), 3-540-73007-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73007-1.

Sachs:2010:ADM

[SPG10] Kai Sachs, Ilia Petrov, and Pablo Guerrero, editors. From Active Data Management to Event-Based Systems and More: Papers in Honor of Alejandro Buchmann on the Occasion of His 60th Birthday, volume 6462 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17225-3 (print), 3-642-17226-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17226-7.

Sawyer:2007:REF

[SPH07] Pete Sawyer, Barbara Paech, and Patrick Heymans, editors. Requirements Engineering: Foundation for Software Quality: 13th International Working Conference, REFSQ 2007, Trondheim, Norway, June 11–12, 2007. Proceedings, volume 4542 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73030-3 (print), 3-540-73031-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73031-6.

Spirakis:1995:AET

[Spi95] P. G. Spirakis, editor. Algorithms — ESA ’95: Third Annual European Symposium, Corfu, Greece, September 25–27, 1995: proceedings, volume 979 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60313-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0979.
htm;   http://www.springerlink.
com/content/978-3-540-
60313-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=979.

Sgall:2001:MFC

[SPK01] Jiri Sgall, Ales Pultr, and Peter Kolman, editors. Mathematical foundations of computer science 2001: 26th International Symposium, MFCS 2001, Marianske Lazne, Czech Republic, August 27–31, 2001: Proceedings, volume 2136 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42496-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2136; QA267 .M38 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2136.htm;
http://www.springerlink.
com/content/978-3-540-
42496-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2136.

Shen:2010:WAI

[SPÖ^{+}10] Heng Tao Shen, Jian Pei, M. Tamer Özsu, Lei Zou, Jiaheng Lu, Tok-Wang Ling, Ge Yu, Yi Zhuang, and Jie Shao, editors. Web-Age Information Management: WAIM 2010 International Workshops: IWGD 2010, XMLDM 2010, WCMT 2010, Jiuzhaigou Valley, China, July 15–17, 2010 Revised Selected Papers, volume 6185 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16719-5 (print), 3-642-16720-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16720-1.

Soille:2011:MMA

[SPO11] Pierre Soille, Martino Pesaresi, and Georgios K.

Ouzounis, editors. Mathematical Morphology and Its Applications to Image and Signal Processing: 10th International Symposium, ISMM 2011, Verbania-Intra, Italy, July 6–8, 2011. Proceedings, volume 6671 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21568-8 (print), 3-642-21569-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21569-8.

Sichman:2008:COI

[SPON08] Jaime Simão Sichman, Julian Padget, Sascha Ossowski, and Pablo Noriega, editors. Coordination, Organizations, Institutions, and Norms in Agent Systems III: COIN 2007 International Workshops COIN@AAMAS 2007, Honolulu, HI, USA, May 14, 2007 COIN@MALLOW 2007, Durham, UK, September 3–4, 2007 Revised Selected Papers, volume 4870 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79002-0 (print), 3-540-79003-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79003-7.

Stuckenschmidt:2009:MOC

[SPS09] Heiner Stuckenschmidt, Christine Parent, and Stefano Spaccapietra, editors. Modular Ontologies: Concepts, Theories and Techniques for Knowledge Modularization, volume 5445 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01906-4 (print), 3-642-01907-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01907-4.

Sole-Pareta:2004:QSE

[SPSV^{+}04] Josep Solé-Pareta, Michael Smirnov, Piet Van Mieghem, Jordi Domingo-Pascual, Edmundo Monteiro, Peter Reichl, Burkhard Stiller, and Richard J. Gibbens, editors. Quality of Service in the Emerging Networking Panorama: Fifth International Workshop on Quality of Future Internet Services, QofIS 2004, and First Workshop on Quality of Service Routing, WQoSR 2004, and Fourth International Workshop on Internet Charging and QoS Technology, ICQT 2004, Barcelona, Catalonia, Spain, September 29–October 1, 2004: proceedings, volume 3266 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23238-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 C67 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3266.
htm;   http://www.springerlink.
com/content/978-3-540-
23238-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3266;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101045.

Spaccapietra:2008:JDSb

[SPT^{+}08] Stefano Spaccapietra, Jeff Z. Pan, Philippe Thiran, Terry Halpin, Steffen Staab, Vojtech Svatek, Pavel Shvaiko, and John Roddick, editors. Journal on Data Semantics XI, volume 5383 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92147-8 (print), 3-540-92148-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92148-6.

Stuller:2000:CID

[ŠPTM00] Julius Štuller, Jaroslav Pokorný, Bernhard Thalheim, and Yoshifumi Masunaga, editors. Current issues in databases and information systems: East-European Conference on Advances in Databases and Information Systems held jointly with International Conference on Database Systems for Advanced Applications, ADBIS-DASFAA 2000, Prague, Czech Republic, September 5–9, 2000: Proceedings, volume 1884 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67977-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 A252 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1884.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67977-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1884.

Schroder-Preikschat:1990:PDO

[SPZ90] W. Schroder-Preikschat and W. Zimmer, editors. Progress in distributed operating systems and distributed systems management: European workshop, Berlin, FRG, April 18–19, 1989: proceedings, volume 433 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52609-9, 0-387-52609-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 P77 1990. DM37.50. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0433.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        52609-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=433.

Proceedings of the First European Workshop on Progress in Distributed Operating Systems and Distributed Systems Management, organized by GMD — German National Research Center for Computer Science.

Soriano:2010:ICS

[SQL10] Miguel Soriano, Sihan Qing, and Javier López, editors. Information and Communications Security: 12th International Conference, ICICS 2010, Barcelona, Spain, December 15–17, 2010. Proceedings, volume 6476 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17649-6 (print), 3-642-17650-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17650-0.

Squazzoni:2009:EAC

[Squ09] Flaminio Squazzoni, editor. Epistemological Aspects of Computer Simulation in the Social Sciences: Second International Workshop, EPOS 2006, Brescia, Italy, October 5–6, 2006, Revised Selected and Invited Papers, volume 5466 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01108-X (print), 3-642-01109-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01109-2.

Schmorrow:2007:FAC

[SR07] Dylan D. Schmorrow and Leah M. Reeves, editors. Foundations of Augmented Cognition: Third International Conference, FAC 2007, Held as Part of HCI International 2007, Beijing, China, July 22–27, 2007. Proceedings, volume 4565 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73215-2 (print), 3-540-73216-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73216-7.

Spring:2011:PAM

[SR11] Neil Spring and George F. Riley, editors. Passive and Active Measurement: 12th International Conference, PAM 2011, Atlanta, GA, USA, March 20–22, 2011. Proceedings, volume 6579 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19259-9 (print), 3-642-19260-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19260-9.

Samarati:2004:CSE

[SRG04] Pierangela Samarati, Peter Ryan, and Dieter Gollmann, editors. Computer Security— ESORICS 2004: 9th European symposium on research in computer security, Sophia Antipolis, France, September 13–15, 2004: Proceedings, volume 3193 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22987-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E98 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3193.
htm;   http://www.springerlink.
com/content/978-3-540-
22987-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3193;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100085.

Su:2012:IRAa

[SRL12a] Chun-Yi Su, Subhash Rakheja, and Honghai Liu, editors. Intelligent robotics and applications: 5th International Conference, ICIRA 2012, Montréal, Canada, October 3–5, 2012, Proceedings. Part I, volume 7506 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33508-X, 3-642-33509-8 (e-book). LCCN TJ210.3 .I23 2012. URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33508-2;   http://www.
                                        springerlink.com/content/
                                        978-3-642-33509-9.

Su:2012:IRAb

[SRL12b] Chun-Yi Su, Subhash Rakheja, and Honghai Liu, editors. Intelligent Robotics and Applications: 5th International Conference, ICIRA 2012, Montreal, Canada, October 3–5, 2012, Proceedings, Part II, volume 7507 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33514-4 (print), 3-642-33515-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33515-0.

Su:2012:IRAc

[SRL12c] Chun-Yi Su, Subhash Rakheja, and Honghai Liu, editors. Intelligent robotics and applications: 5th International Conference, ICIRA 2012, Montréal, QC, Canada, October 3–5, 2012, Proceedings. Part III, volume 7508 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33502-0, 3-642-33503-9 (e-book). LCCN TJ210.3 .I25 2012. URL http:

//www.springerlink.
com/content/978-3-642-
33502-0;   http://www.
springerlink.com/content/
978-3-642-33503-7.

Stilla:2011:PIA

[SRM^{+}11] Uwe Stilla, Franz Rottensteiner, Helmut Mayer, Boris Jutzi, and Matthias Butenuth, editors. Photogrammetric Image Analysis: ISPRS Conference, PIA 2011, Munich, Germany, October 5–7, 2011. Proceedings, volume 6952 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24392-4 (print), 3-642-24393-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24393-6.

Sanfeliu:2003:PPR

[SRS03] Alberto Sanfeliu and José Ruiz-Shulcloper, editors. Progress in Pattern Recognition, Speech and Image Analysis: 8th Iberoamerican Congress on pattern recognition, CIARP 2003, Havana, Cuba, November 26– 29, 2003: Proceedings, volume 2905 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20590-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2905.htm;
 http://www.springerlink.
com/content/978-3-540-
20590-6;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2905;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94613.

Stiller:2006:PPI

[SRT06] Burkhard Stiller, Peter Reichl, and Bruno Tuffin, editors. Performability Has its Price: 5th International Workshop on Internet Charging and QoS Technologies, ICQT 2006, St. Malo, France, June 27, 2006. Proceedings, volume 4033 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35456-5 (print), 3-540-35457-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35457-4.

Singh:1998:IAI

[SRW98] Munindar P. (Munindar Paul) Singh, Anand Rao, and Michael J. Wooldridge, editors. Intelligent agents IV: agent theories, architectures, and languages: 4th International Workshop, ATAL’97, Providence, Rhode Island, USA, July 24–26, 1997 proceedings, volume 1365 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64162-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58A83 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1365.htm;
 http://www.springerlink.
com/content/978-3-540-
64162-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1365.

Srinathan:2007:PCI

[SRY07] K. Srinathan, C. Pandu Rangan, and Moti Yung, editors. Progress in Cryptology — INDOCRYPT 2007: 8th International Conference on Cryptology in India, Chennai, India, December 9– 13, 2007. Proceedings, volume 4859 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77025-9 (print), 3-540-77026-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77026-8.

Specker:1976:KES

[SS76] Ernst Specker and Volker Strassen, editors. Komplexität von Entscheidungsproblemen: ein Seminar. (German) [Complexity of Decision Problems: A Seminar], volume 43 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1976. CODEN LNCSD9. ISBN 0-387-07805-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.65 .K65. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0043.htm;
http://www.springerlink.
com/content/978-0-387-
07805-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=43.

Salomaa:1977:ALP

[SS77] Arto Salomaa and Magnus Steinby, editors. Automata, languages and programming: fourth colloquium, University of Turku, Finland, July 18– 22, 1977, volume 52 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CO-DEN LNCSD9. ISBN 0-387-08342-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.52.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0052.
htm;   http://www.springerlink.
com/content/978-0-387-
08342-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=52.

Salton:1983:RDI

[SS83] Gerard Salton and Hans-Jochen Schneider, editors. Research and development in information retrieval: proceedings, Berlin, May 18– 20, 1982, volume 146 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-11978-7 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z699.A1 R47 1982. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0146.htm;
 http://www.springerlink.
com/content/978-0-387-
11978-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=146.

Schmidt-Schauss:1989:CAO

[SS89] M. (Manfred) Schmidt-Schauss. Computational aspects of an order-sorted logic with term declarations, volume 395 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51705-7 (USA) (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 171 pp. LCCN QA9 .S343 1989. DM33.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0395.
htm;   http://www.springerlink.
com/content/978-0-387-
51705-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=395.

Simons:1990:FTD

[SS90] B. Simons and Alfred Z.

Spector. Fault-tolerant distributed computing, volume 448 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 0-387-97385-0. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 208 pp. LCCN QA76.9.D5 F38 1990. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0448.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        97385-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=448.

“Bibliography for faulttolerant distributed computing”: p. [274]-298.

Schmidt:1991:NGI

[SS91] Joachim W. Schmidt and A. A. Stognii, editors. Next generation information system technology: proceedings, volume 504 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54141-1 (Berlin), 0-387-54141-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I565 1990. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0504.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54141-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=504.

Schmidt:1994:ATS

[SS94] E. M. Schmidt and Sven Skyum, editors. Algorithm theory — SWAT ’94: 4th Scandinavian Workshop on Algorithm Theory, Aarhus, Denmark, July 6–8, 1994, proceedings, volume 824 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58218-5 (Berlin), 0-387-58218-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .S28 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0824.htm;
 http://www.springerlink.
com/content/978-3-540-
58218-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=824.

Stadler:1999:ATN

[SS99] Rolf Stadler and Burkhard Stiller, editors. Active technologies for network and service management: 10th IFIP/IEEE International Workshop on Distributed Systems: Operations and Management, DSOM’99, Zürich, Switzerland, October 11– 13, 1999: Proceedings, volume 1700 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66598-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I33835 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1700.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66598-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1700.

Schiele:2001:CVS

[SS01] Bernard Schiele and Gerhard Sagerer, editors. Computer vision systems: Second International Workshop, ICVS 2001, Vancouver, Canada, July 7–8, 2001: Proceedings, volume 2095 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42285-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .I58 2001; QA267.A1 L43 no.2095. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2095.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42285-3;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=2095.

Stary:2004:UCI

[SS04] Christian Stary and Constantine Stephanidis, editors. User-Centered Interaction Paradigms for Universal Access in the Information Society: 8th ERCIM Workshop on User Interfaces for All, Vienna, Austria, June 28–29, 2004: revised Selected Papers, volume 3196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23375-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.U83 E73 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3196.
htm;   http://www.springerlink.
com/content/978-3-540-
23375-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3196;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95185.

Sahin:2005:SRS

[ŞS05a] Erol Şahin and William M.

Spears, editors. Swarm robotics: SAB 2004 International Workshop, Santa Monica, CA, USA, July 17, 2004. Revised selected papers, volume 3342 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24296-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ210.3. URL

http://www.springerlink.
com/content/978-3-540-
24296-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3336;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105069.

Schonwalder:2005:ANI

[SS05b] Jürgen Schönwälder and Joan Serrat, editors. Ambient networks: 16th IFIP/IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2005, Barcelona, Spain, October 24–26, 2005: proceedings, volume 3775 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29388-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I33835 2005.

URL http://www.springerlink.
com/content/978-3-540-
29388-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3775.

Schwentick:2006:DTI

[SS06a] Thomas Schwentick and Dan Suciu, editors. Database Theory — ICDT 2007: 11th International Conference, Barcelona, Spain, January 10–12, 2007. Proceedings, volume 4353 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69269-X (print), 3-540-69270-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69270-6.

Shi:2006:ACM

[SS06b] Zhong-Zhi Shi and Ramakoti Sadananda, editors. Agent Computing and Multi-Agent Systems: 9th Pacific Rim International Workshop on Multi-Agents, PRIMA 2006, Guilin, China, August 7–8, 2006. Proceedings, volume 4088 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36707-1 (print), 3-540-36860-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        36860-1.

Staab:2006:MKW

[SS06c] Steffen Staab and Vojtěch Svátek, editors. Managing Knowledge in a World of Networks: 15th International Conference, EKAW 2006, Poděbrady, Czech Republic, October 2–6, 2006. Proceedings, volume 4248 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46363-1 (print), 3-540-46365-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46365-8.

Stavrakakis:2006:ACS

[SS06d] Ioannis Stavrakakis and Michael Smirnov, editors. Autonomic communication: Second international IFIP workshop, WAC 2005, Athens, Greece, October 2–5, 2005, revised selected papers, volume 3854 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32992-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32992-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3854.

Stock:2006:RAI

[SS06e] Oliviero Stock and Marco Schaerf, editors. Reasoning, Action and Interaction in AI Theories and Systems: Essays Dedicated to Luigia Carlucci Aiello, volume 4155 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37901-0 (print), 3-540-37902-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37902-7.

Sachse:2007:FIM

[SS07a] Frank B. Sachse and Gunnar Seemann, editors. Functional Imaging and Modeling of the Heart: 4th International Conference, FIHM 2007, Salt Lake City, UT, USA, June 7–9, 2007. Proceedings, volume 4466 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72906-2 (print), 3-540-72907-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72907-5.

Smith:2007:HIMb

[SS07b] Michael J. Smith and Gavriel Salvendy, editors. Human Interface and the Management of Information. Interacting in Information Environments: Symposium on Human Interface 2007, Held as Part of HCI International 2007, Beijing, China, July 22– 27, 2007, Proceedings, Part II, volume 4558 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73353-1 (print), 3-540-73354-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
73354-6.

Smith:2007:HIMa

[SS07c] Michael J. Smith and Gavriel Salvendy, editors. Human Interface and the Management of Information. Methods, Techniques and Tools in Information Design: Symposium on Human Interface 2007, Held as Part of HCI International 2007, Beijing, China, July 22–27, 2007, Proceedings Part I, volume 4557 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73344-2 (print), 3-540-73345-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73345-4.

Sadri:2008:CLM

[SS08a] Fariba Sadri and Ken Satoh, editors. Computational Logic in Multi-Agent Systems: 8th International Workshop, CLIMA VIII, Porto, Portugal, September 10– 11, 2007. Revised Selected and Invited Papers, volume 5056 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88832-2 (print), 3-540-88833-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88833-8.

Spierling:2008:ISF

[SS08b] Ulrike Spierling and Nicolas Szilas, editors. Interactive Storytelling: First Joint International Conference on Interactive Digital Storytelling, ICIDS 2008 Erfurt, Germany, November 26– 29, 2008 Proceedings, volume 5334 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89424-1 (print), 3-540-89454-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89454-4.

Salvendy:2009:HIM

[SS09a] Gavriel Salvendy and Michael J. Smith, editors. Human Interface and the Management of Information. Information and Interaction: Symposium on Human Interface 2009, Held as part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009, Proceedings, Part II, volume 5618 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02558-7 (print), 3-642-02559-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02559-4.

Schurr:2009:MDE

[SS09b] Andy Schürr and Bran Selic, editors. Model Driven Engineering Languages and Systems: 12th International Conference, MODELS 2009, Denver, CO, USA, October 4–9, 2009. Proceedings, volume 5795 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04424-7 (print), 3-642-04425-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        04425-0.

Smith:2009:HIM

[SS09c] Michael J. Smith and Gavriel Salvendy, editors. Human Interface and the Management of Information. Designing Information Environments: Symposium on Human Interface 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009, Proceedings, Part I, volume 5617 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02555-2 (print), 3-642-02556-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02556-3.

Stevens:2009:ECI

[SS09d] Scott M. Stevens and Shirley J. Saldamarco, editors. Entertainment Computing — ICEC 2008: 7th International Conference, Pittsburgh, PA, USA, September 25– 27, 2008. Proceedings, volume 5309 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-89221-4 (print), 3-540-89222-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89222-9.

Strichman:2010:TAS

[SS10] Ofer Strichman and Stefan Szeider, editors. Theory and Applications of Satisfiability Testing – SAT 2010: 13th International Conference, SAT 2010, Edinburgh, UK, July 11–14, 2010. Proceedings, volume 6175 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14185-4 (print), 3-642-14186-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14186-7.

Sakallah:2011:TAS

[SS11a] Karem A. Sakallah and Laurent Simon, editors. Theory and Applications of Satisfiability Testing — SAT 2011: 14th International Conference, SAT 2011, Ann Arbor, MI, USA, June 19– 22, 2011. Proceedings, volume 6695 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21580-7 (print), 3-642-21581-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21581-0.

Salaun:2011:FMI

[SS11b] Gwen Salaün and Bernhard Schätz, editors. Formal Methods for Industrial Critical Systems: 16th International Workshop, FMICS 2011, Trento, Italy, August 29–30, 2011. Proceedings, volume 6959 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24430-0 (print), 3-642-24431-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24431-5.

Salvendy:2011:HIM

[SS11c] Gavriel Salvendy and Michael J. Smith, editors. Human Interface and the Management of Information. Interacting with Information: Symposium on Human Interface 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part II, volume 6772 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21668-4 (print), 3-642-21669-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21669-5.

Smith:2011:HIM

[SS11d] Michael J. Smith and Gavriel Salvendy, editors. Human Interface and the Management of Information. Interacting with Information: Symposium on Human Interface 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part I, volume 6771 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21792-3 (print), 3-642-21793-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21793-7.

Singh:2005:TICa

[SSAP05a] Sameer Singh, Maneesha Singh, Chid Apte, and Petra Perner, editors. Third International Conference on Advances in Pattern Recognition: ICAPR 2005, Bath, UK, August 22–25, 2005: proceedings, volume 3686 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28757-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 I566 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        28757-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3686.

Singh:2005:TICb

[SSAP05b] Sameer Singh, Maneesha Singh, Chid Apte, and Petra Perner, editors. Third International Conference on Advances in Pattern Recognition: ICAPR 2005, Bath, UK, August 22–25, 2005: proceedings, volume 3687 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28833-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 I566 2005.

URL http://www.springerlink.
com/content/978-3-540-
28833-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3687.

Steinholtz:1990:AIS

[SSB90] B. Steinholtz, Arne Sølvberg, and L. Bergman, editors. Advanced information systems engineering: Second Nordic Conference CAiSE ’90, Stockholm, Sweden, May 8–10, 1990: proceedings, volume 436 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52625-0 (Berlin), 0-387-52625-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .N67 1990. DM57.50. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0436.
htm;   http://www.springerlink.
com/content/978-3-540-
52625-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=436.

Sheth:2008:SWI

[SSD^{+}08] Amit Sheth, Steffen Staab, Mike Dean, Massimo Paolucci, Diana Maynard, Timothy Finin, and Krishnaprasad Thirunarayan, editors. The Semantic Web — ISWC 2008: 7th International Semantic Web Conference, ISWC 2008, Karlsruhe, Germany, October 26– 30, 2008. Proceedings, volume 5318 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88563-3 (print), 3-540-88564-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88564-1.

Serna:2010:ARC

[SSJR10] Maria Serna, Ronen Shaltiel, Klaus Jansen, and José Rolim, editors. Approximation, Randomization, and Combinatorial Optimization. Algorithms and Techniques: 13th International Workshop, AP-PROX 2010, and 14th International Workshop, RANDOM 2010, Barcelona, Spain, September 1– 3, 2010. Proceedings, volume 6302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15368-2 (print), 3-642-15369-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15369-3.

Stiller:2002:QPQ

[SSKR02] Burkhard Stiller, Michael Smirnow, Martin Karsten, and Peter Reichl, editors. From QoS provisioning to QoS charging: third COST 263 International Workshop on Quality of Future Internet Services, QofIS 2002 and Second International Workshop on Internet Charging and QoS Technologies, ICQT 2002, Zürich, Switzerland, October 16–18, 2002: Proceedings, volume 2511 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44356-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 C67 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2511.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44356-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2511.

Solanki:2008:IHI

[SSM08] Kaushal Solanki, Kenneth Sullivan, and Upamanyu Madhow, editors. Information Hiding: 10th International Workshop, IH 2008, Santa Barbara, CA, USA, May 19–21, 2008, Revised Selected Papers, volume 5284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88960-4 (print), 3-540-88961-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88961-8.

Schulzrinne:2008:PSA

[SSN08]  Henning  Schulzrinne,

Radu State, and Saverio Niccolini, editors. Principles, Systems and Applications of IP Telecommunications. Services and Security for Next Generation Networks: Second International Conference, IPTComm 2008, Heidelberg, Germany, July 1–2, 2008. Revised Selected Papers, volume 5310 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89053-X (print), 3-540-89054-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89054-6.

Salem:2011:IRT

[SSO^{+}11] Mohamed Vall Mohamed Salem, Khaled Shaalan, Farhad Oroumchian, Azadeh Shakery, and Halim Khelalfa, editors. Information Retrieval Technology: 7th Asia Information Retrieval Societies Conference, AIRS 2011, Dubai, United Arab Emirates, December 18–20, 2011. Proceedings, volume 7097 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25630-9 (print), 3-642-25631-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25631-8.

Schek:1998:ADT

[SSRA98] Hans-Jörg Schek, Felix Saltor, Isidro Ramos, and Gustavo Alonso, editors. Advances in database technology: EDBT ’98: 6th International Conference on Extending Database Technology, Valencia, Spain, March 23–27, 1998: Proceedings, volume 1377 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64264-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3I55844 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1377.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64264-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1377.

Saake:2000:TDD

[SST00] Gunter Saake, Kerstin Schwarz, and Can Türker, editors. Transactions and database dynamics: 8th International Workshop on Foundations of Models and Languages for Data and Objects, Dagstuhl Castle, Germany, September 27–30, 1999: Selected Papers, volume 1773 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67201-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1773.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1773.
htm;   http://www.springerlink.
com/content/978-3-540-
67201-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1773.

Stachniss:2012:SCV

[SSU12] Cyrill Stachniss, Kerstin Schill, and David Uttal, editors. Spatial Cognition VIII: International Conference, Spatial Cognition 2012, Kloster Seeon, Germany, August 31 — September 3, 2012. Proceedings, volume 7463 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32731-1 (print), 3-642-32732-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32732-2.

Steels:1994:FKA

[SSv94] Luc Steels, Guus Schreiber, and Walter van de Velde, editors. A future for knowledge acquisition: 8th European Knowledge Acquisition Workshop, EKAW ’94, Hoegaarden, Belgium, September 26–29, 1994: proceedings, volume 867 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58487-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E36 1994.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0867.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        58487-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=867.

Sakama:2012:DAL

[SSVW12] Chiaki Sakama, Sebastian Sardina, Wamberto Vasconcelos, and Michael Winikoff, editors. Declarative Agent Languages and Technologies IX: 9th International Workshop, DALT 2011, Taipei, Taiwan, May 3, 2011, Revised Selected and Invited Papers, volume 7169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29112-0 (print), 3-642-29113-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29113-5.

Schmalhofer:1992:CKE

[SSW92] F. Schmalhofer, Gerhard Strube, and Thomas Wetter, editors. Contemporary knowledge engineering and cognition: first joint workshop, Kaiserslautern, Germany, February 21– 22, 1991, proceedings, volume 622 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55711-3 (Berlin), 0-387-55711-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 C667 1992. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t0622.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55711-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=622.

Sahin:2007:SRS

[ŞSW07] Erol Şahin, William M. Spears, and Alan F. T. Winfield, editors. Swarm Robotics: Second International Workshop, SAB 2006, Rome, Italy, September 30–October 1, 2006, Revised Selected Papers, volume 4433 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71540-1 (print), 3-540-71541-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71541-2.

Szmuc:2012:ASE

[SSZ12] Tomasz Szmuc, Marcin Szpyrka, and Jaroslav Zendulka, editors. Advances in Software Engineering Techniques: 4th IFIP TC 2 Central and East European Conference on Software Engineering Techniques, CEE-SET 2009, Krakow, Poland, October 12–14, 2009. Revised Selected Papers, volume 7054 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28037-4 (print), 3-642-28038-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28038-2.

Snyers:1987:LDL

[ST87] Dominique Snyers and Andre Thayse. From logic design to logic programming: theorem proving techniques and Pfunctions, volume 271 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18217-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). iv + 125 pp. LCCN QA76.9.A96 S621 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0271.htm;
 http://www.springerlink.
com/content/978-0-387-
18217-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=271.

Sannella:1988:RTD

[ST88] D. Sannella and Andrzej Tarlecki, editors. Recent trends in data type specification: 5th Workshop on Specification of Abstract Data Types: selected papers, Gullane, Scotland, September 1– 4, 1987, volume 332 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-50325-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A23 W671 1987.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0332.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        50325-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=332.

Sanchez:1996:TEH

[ST96] Eduardo Sanchez and Marco Tomassini, editors. Towards evolvable hardware: the evolutionary engineering approach, volume 1062 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61093-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7888.4 .T69 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1062.htm;
http://www.springerlink.
com/content/978-3-540-
61093-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1062.

Schewe:2000:FIK

[ST00] Klaus-Dieter Schewe and B. Thalheim, editors. Foundations of information and knowledge systems: First International Symposium, FoIKS 2000, Burg, Germany, February 2000: Proceedings, volume 1762 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67100-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1762.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1762.
htm;   http://www.springerlink.
com/content/978-3-540-
67100-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1762.

Stinson:2001:SAC

[ST01] Douglas R. Stinson and Stafford Tavares, editors. Selected areas in cryptography: 7th annual International Workshop, SAC 2000, Waterloo, Ontario, Canada, August 14–15, 2000: Proceedings, volume 2012 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42069-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 S22 2000; QA267.A1 L43 no.2012. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2012.htm;
 http://www.springerlink.
com/content/978-3-540-
42069-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2012.

Shafazand:2002:EII

[ST02] M. Hassan Shafazand and A. Min Tjoa, editors. EurAsia-ICT 2002: Information and Communication Technology: First EurAsian Conference Shiraz, Iran, October 29–31, 2002. Proceedings, volume 2510 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00028-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T58.5 .E818 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2510.htm;
http://www.springerlink.
com/content/978-3-540-
00028-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2510.

Sokolsky:2007:RVI

[ST07] Oleg Sokolsky and Serdar Taşıran, editors. Runtime Verification: 7th International Workshop, RV 2007, Vancover, Canada, March 13, 2007, Revised Selected Papers, volume 4839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77394-0 (print), 3-540-77395-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77395-5.

Schewe:2008:SDK

[ST08] Klaus-Dieter Schewe and Bernhard Thalheim, editors. Semantics in Data and Knowledge Bases: Third International Workshop, SDKB 2008, Nantes, France, March 29, 2008, Revised Selected Papers, volume 4925 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88593-5 (print), 3-540-88594-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88594-8.

Schewe:2011:SDK

[ST11] Klaus-Dieter Schewe and Bernhard Thalheim, editors. Semantics in Data and Knowledge Bases: 4th International Workshops, SDKB 2010, Bordeaux, France, July 5, 2010, Revised Selected Papers, volume 6834 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23440-2 (print), 3-642-23441-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23441-5.

Schrijvers:2012:FLP

[ST12a] Tom Schrijvers and Peter Thiemann, editors. Functional and Logic Programming: 11th International Symposium, FLOPS 2012, Kobe, Japan, May 23–25, 2012. Proceedings, volume 7294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29821-4 (print), 3-642-29822-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29822-6.

Schwenker:2012:PSL

[ST12b]  Friedhelm  Schwenker

and Edmondo Trentin, editors. Partially Supervised Learning: First IAPR TC3 Workshop, PSL 2011, Ulm, Germany, September 15–16, 2011, Revised Selected Papers, volume 7081 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28257-1 (print), 3-642-28258-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28258-4.

Stefanovic:2012:DCM

[ST12c] Darko Stefanovic and Andrew Turberfield, editors. DNA Computing and Molecular Programming: 18th International Conference, DNA 18, Aarhus, Denmark, August 14–17, 2012. Proceedings, volume 7433 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32207-7 (print), 3-642-32208-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32208-2.

Staunstrup:1982:PSP

[Sta82] J. Staunstrup, editor. Program specification: proceedings of a workshop, Aarhus, Denmark, August 1981, volume 134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CODEN LNCSD9. ISBN 0-387-11490-4 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75119. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0134.
                                        htm;   http://www.springerlink.

com/content/978-0-387-
11490-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=134.

Staples:1995:ACI

[Sta95] John Staples, editor. Algorithms and computations: 6th international symposium, ISAAC ’95, Cairns, Australia, December 4–6, 1995: proceedings, volume 1004 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60573-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 I73 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1004.
htm;   http://www.springerlink.
com/content/978-3-540-
60573-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1004.

Staab:1999:GKE

[Sta99] Steffen Staab. Grading Knowledge: Extracting Degree Information from Texts, volume 1744 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66934-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 187 pp. LCCN QA76.9.N38 S78 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1744.htm;
 http://www.springerlink.
com/content/978-3-540-
66934-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1744.

Si:2011:ISF

[STA^{+}11] Mei Si, David Thue, Elisabeth André, James C. Lester, Joshua Tanenbaum, and Veronica Zammitto, editors. Interactive Storytelling: Fourth International Conference on Interactive Digital Storytelling, ICIDS 2011, Vancouver, Canada, November 28 — 1 December, 2011. Proceedings, volume 7069 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25288-5 (print), 3-642-25289-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25289-1.

Sanfeliu:2004:PPR

[STC04] Alberto Sanfeliu, Jose Francisco Martínez Trinidad, and Jesús Ariel Carrasco Ochoa, editors. Progress in pattern recognition, image analysis and applications: 9th Iberoamerican congress on pattern recognition, CIARP 2004, Puebla, Mexico, October 26–29, 2004, proceedings, volume 3287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23527-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
23527-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3287;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101756.

Sloot:2002:CSIb

[STDH02a] P. M. A. Sloot, C. J. Kenneth Tan, J. J. Dongarra, and A. G. Hoekstra, editors. Computational Science—ICCS 2002: International Conference, Amsterdam, The Netherlands, April 21– 24, 2002. Proceedings, Part II, volume 2330 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43593-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2002 pt.1-3 (2002). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2330.htm;
 http://www.springerlink.
com/content/978-3-540-
43593-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2330.

Sloot:2002:CSIc

[STDH02b] P. M. A. Sloot, C. J. Kenneth Tan, J. J. Dongarra, and A. G. Hoekstra, editors. Computational Science—ICCS 2002: International Conference, Amsterdam, The Netherlands, April 21– 24, 2002. Proceedings, Part III, volume 2331 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43594-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2002 pt.1-3 (2002). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2331.htm;
 http://www.springerlink.
com/content/978-3-540-
43594-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2331.

Sloot:2002:CSIa

[STDH02c] Peter M. A. Sloot, C. J. Kenneth Tan, Jack J. Dongarra, and Alfons G. Hoekstra, editors. Computational Science—ICCS 2002: International Conference, Amsterdam, The Netherlands, April 21– 24, 2002. Proceedings, Part I, volume 2329 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43591-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2002 pt.1-3 (2002). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2329.htm;
 http://www.springerlink.
com/content/978-3-540-
43591-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2329.

Steinmetz:1994:MAT

[Ste94] Ralf Steinmetz, editor. Multimedia: advanced teleservices and highspeed communication architectures: second international workshop, IWACA ’94, Heidelberg, Germany, September 26–28, 1994: proceedings, volume 868 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58494-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I93 1994. DM86.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0868.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58494-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=868.

Steffens:1995:MTL

[Ste95] Petra Steffens, editor. Machine translation and the lexicon: third International EAMT Workshop, Heidelberg, Germany, April 26–28, 1993: proceedings, volume 898 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 0-387-59040-4 (USA), 3-540-59040-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN P309 .I57 1993. DM62.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0898.htm;
 http://www.springerlink.
com/content/978-0-387-
59040-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=898.

Steffen:1998:TAC

[Ste98] Bernhard Steffen, editor. Tools and algorithms for the construction and analysis of systems: 4th International Conference, TACAS ’98, held as part of the joint European Conferences on Theory and Practice of Software, ETAPS ’98, Lisbon, Portugal, March/April 1998: Proceedings, volume 1384 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64356-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88T33 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1384.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        64356-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1384.

Stern:1999:ACE

[Ste99] J. Stern, editor. Advances in cryptology — EUROCRYPT ’99: International Conference on the theory and application of cryptographic techniques, Prague, Czech Republic, May 2–6, 1999: Proceedings, volume 1592 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65889-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 E964 1999.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1592.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65889-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1592.

Steinhofel:2001:SAF

[Ste01] Kathleen Steinhöfel, editor. Stochastic Algorithms: Foundations and Applications: International Symposium, SAGA 2001, Berlin, Germany, December 13– 14, 2001: Proceedings, volume 2264 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43025-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .S24 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2264.htm;
 http://www.springerlink.
com/content/978-3-540-
43025-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2264.

Stephanidis:2005:UAH

[Ste05] Constantine Stephanidis.

Universal access in health telematics: a design code of practice, volume 3041 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26167-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26167-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=3041;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136257.

Stenstrom:2007:THP

[Ste07a] Per Stenström, editor. Transactions on High-Performance Embedded Architectures and Compilers I, volume 4050 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71527-4 (print), 3-540-71528-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71528-3.

Stephanidis:2007:UAHb

[Ste07b] Constantine Stephanidis, editor. Universal Access in Human-Computer Interaction. Ambient Interaction: 4th International Conference on Universal Access in Human-Computer Interaction, UAHCI 2007 Held as Part of HCI International 2007 Beijing, China, July 22–27, 2007 Proceedings, Part II, volume 4555 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73280-2 (print), 3-540-73281-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73281-5.

Stephanidis:2007:UAHc

[Ste07c] Constantine Stephanidis, editor. Universal Access in Human-Computer Interaction. Applications and Services: 4th International Conference on Universal Access in Human-Computer Interaction, UAHCI 2007 Held as Part of HCI International 2007 Beijing, China, July 22–27, 2007 Proceedings, Part III, volume 4556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73282-9 (print), 3-540-73283-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73283-9.

Stephanidis:2007:UAHa

[Ste07d] Constantine Stephanidis, editor. Universal Acess in Human Computer Interaction. Coping with Diversity: 4th International Conference on Universal Access in Human-Computer Interaction, UAHCI 2007, Held as Part of HCI International 2007, Beijing, China, July 22– 27, 2007, Proceedings, Part I, volume 4554 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73278-0 (print), 3-540-73279-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73279-2.

Stenstrom:2009:THP

[Ste09a] Per Stenström, editor. Transactions on High-Performance Embedded Architectures and Compilers II, volume 5470 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00903-4 (print), 3-642-00904-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00904-4.

Stephanidis:2009:UAHa

[Ste09b] Constantine Stephanidis, editor. Universal Access in Human-Computer Interaction. Addressing Diversity: 5th International Conference, UAHCI 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009. Proceedings, Part I, volume 5614 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02706-7 (print), 3-642-02707-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02707-9.

Stephanidis:2009:UAHc

[Ste09c] Constantine Stephanidis, editor. Universal Access in Human-Computer Interaction. Applications and Services: 5th International Conference, UAHCI 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009. Proceedings, Part III, volume 5616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02712-1 (print), 3-642-02713-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02713-0.

Stephanidis:2009:UAHb

[Ste09d] Constantine Stephanidis, editor. Universal Access in Human-Computer Interaction. Intelligent and Ubiquitous Interaction Environments: 5th International Conference, UAHCI 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19–24, 2009. Proceedings, Part II, volume 5615 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02709-1 (print), 3-642-02710-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02710-9.

Stenstrom:2011:THPa

[Ste11a] Per Stenström, editor. Transactions on High-Performance Embedded Architectures and Compilers III, volume 6590 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19447-8 (print), 3-642-19448-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19448-1.

Stenstrom:2011:THPb

[Ste11b] Per Stenström, editor.

Transactions on High-Performance Embedded Architectures and Compilers IV, volume 6760 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24567-6 (print), 3-642-24568-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24568-8.

Stephanidis:2011:UAHd

[Ste11c] Constantine Stephanidis, editor. Universal Access in Human-Computer Interaction. Applications and Services: 6th International Conference, UAHCI 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part IV, volume 6768 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21656-0 (print), 3-642-21657-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21657-2.

Stephanidis:2011:UAHc

[Ste11d] Constantine Stephanidis, editor. Universal Access in Human-Computer Interaction. Context Diversity: 6th International Conference, UAHCI 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part III, volume 6767 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21665-X (print), 3-642-21666-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21666-4.

Stephanidis:2011:UAHa

[Ste11e] Constantine Stephanidis, editor. Universal Access in Human-Computer Interaction. Design for All and eInclusion: 6th International Conference, UAHCI 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part I, volume 6765 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21671-4 (print), 3-642-21672-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21672-5.

Stephanidis:2011:UAHb

[Ste11f] Constantine Stephanidis, editor. Universal Access in Human-Computer Interaction. Users Diversity: 6th International Conference, UAHCI 2011, Held as Part of HCI International 2011, Orlando, FL, USA, July 9–14, 2011, Proceedings, Part II, volume 6766 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21662-5 (print), 3-642-21663-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        21663-3.

Steels:2012:CIF

[Ste12] Luc Steels, editor. Computational issues in fluid construction grammar: a new formalism for the representation of lexicons and grammars, volume 7249 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34119-5, 3-642-34120-9 (e-book). ???? pp. LCCN P98 .C66 2012. URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        34120-5.

Stickel:1990:PIC

[Sti90] M. E. Stickel, editor. Proceedings: 10th International Conference on Automated Deduction, Kaiserslautern, FRG, July 24–27, 1990, volume 449 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-52885-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I57 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0449.htm;
 http://www.springerlink.
com/content/978-0-387-
52885-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=449.

Stinson:1994:ACC

[Sti94] Douglas R. Stinson, editor. Advances in Cryptology, CRYPTO ’93: 13th Annual International Cryptology Conference, Santa Barbara, California, USA, August 22–26, 1993: Proceedings, volume 773 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57766-1 (New York), 3-540-57766-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0773.htm;
 http://www.springerlink.
com/content/978-0-387-
57766-1;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=773.

Schmidhuber:2011:AGI

[STL11] Jürgen Schmidhuber, Kristinn R. Thórisson, and Moshe Looks, editors. Artificial General Intelligence: 4th International Conference, AGI 2011, Mountain View, CA, USA, August 3–6, 2011. Proceedings, volume 6830 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22886-0 (print), 3-642-22887-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22887-2.

Su:2005:NLP

[STLK05] Keh-Yih Su, Jun’ichi Tsujii, Jong-Hyeok Lee, and Oi Yee Kwong, editors. Natural language processing — IJC-NLP 2004: first international joint conference, Hainan Island, China, March 22–24, 2004, revised selected papers, volume 3248 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24475-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
24475-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3248.

Stoica:2004:SCS

[Sto04] Ion Stoica. Stateless Core: A Scalable Approach for Quality of Service in the Internet, volume 2979 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21960-9. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 219 pp. LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2979.htm;
 http://www.springerlink.
com/content/978-3-540-
21960-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2979;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/

b97828. Winning thesis of the 2001 ACM Doctoral Dissertation Competition.

Samarati:2010:IST

[STP^{+}10]  Pierangela  Samarati,

Michael Tunstall, Joachim Posegga, Konstantinos Markantonakis, and Damien Sauveron, editors. Information Security Theory and Practices. Security and Privacy of Pervasive Systems and Smart Devices: 4th IFIP WG 11.2 International Workshop, WISTP 2010, Passau, Germany, April 12–14, 2010. Proceedings, volume 6033 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12367-8 (print), 3-642-12368-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12368-9.

Strohmeier:1996:RST

[Str96] Alfred Strohmeier, editor. Reliable software technologies, Ada-Europe ’96: 1996 Ada-Europe International Conference on Reliable Software Technologies, Montreux, Switzerland, June 10–14, 1996: proceedings, volume 1088 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61317-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1088.htm;
 http://www.springerlink.
com/content/978-3-540-
61317-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1088.

Streitz:1999:CBI

[Str99] N. (Norbert) Streitz, editor. Cooperative buildings: integrating information, organizations, and architecture: Second International Workshop, CoBuild ’99, Pittsburgh, USA, October 1- 2, 1999 proceedings, volume 1670 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66596-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN NA2728 .C59 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1670.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        66596-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1670.

Shawe-Taylor:2004:LTA

[STS04] John Shawe-Taylor and Yoram Singer, editors. Learning Theory: 17th Annual Conference on Learning Theory, COLT 2004, Banff, Canada, July 1–4, 2004: Proceedings, volume 3120 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22282-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3120.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22282-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3120;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98522.

Seipel:2004:FIK

[STT04] Dietmar Seipel and José María Turull-Torres, editors. Foundations of Information and Knowledge Systems: Third International Symposium, FoIKS 2004: Wilheminenburg [sic] Castle, Austria, February 17– 20, 2004: Proceedings, volume 2942 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20965-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .F65 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2942.htm;
 http://www.springerlink.
com/content/978-3-540-
20965-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2942;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95342.

Schettini:2011:CCI

[STT11] Raimondo Schettini, Shoji Tominaga, and Alain Trémeau, editors. Computational Color Imaging: Third International Workshop, CCIW 2011, Milan, Italy, April 20– 21, 2011. Proceedings, volume 6626 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20403-1 (print), 3-642-20404-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20404-3.

Sterbenz:2002:ANI

[STTP02] James P. G. Sterbenz, Osamu Takada, Christian Tschudin, and Bernhard Plattner, editors. Active Networks: IFIP-TC6 4th International Working Conference, IWAN 2002, Zürich, Switzerland, December 4–6, 2002. Proceedings, volume 2546 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00223-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I95 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2546.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00223-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2546.

Also available via the

World Wide Web.

Stuttgen:1985:HAP

[Stu85] Heinrich J. Stuttgen.

A hierarchical associative processing system, volume 195 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15652-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 273 pp. LCCN TK7895.M4 S781 1985. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0195.htm;
 http://www.springerlink.
com/content/978-0-387-
15652-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=195.

Studer:1990:NLL

[Stu90] Rudi Studer, editor.

Natural language and logic: International Scientific Symposium, Hamburg, FRG, May 9–11, 1989, volume 459 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 0-387-53082-7 (New York), 3-540-53082-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.N38 I57 1989. DM42.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0459.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        53082-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=459.

Stuckey:2002:LPI

[Stu02] Peter J. Stuckey, editor. Logic Programming: 18th International Conference, ICLP 2002, Copenhagen, Denmark, July 29–August 1, 2002. Proceedings, volume 2401 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43930-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I56a 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2401.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43930-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2401.

Stuckey:2008:PPC

[Stu08] Peter J. Stuckey, editor. Principles and Practice of Constraint Programming: 14th International Conference, CP 2008, Sydney, Australia, September 14–18, 2008. Proceedings, volume 5202 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85957-8 (print), 3-540-85958-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85958-1.

Stutzle:2009:LIO

[Stü09] Thomas Stützle, editor. Learning and Intelligent Optimization: Third International Conference, LION 3, Trento, Italy, January 14–18, 2009. Selected Papers, volume 5851 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-11168-8 (print), 3-642-11169-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11169-3.

Stephens:2007:FGA

[STWS07] Christopher R. Stephens, Marc Toussaint, Darrell Whitley, and Peter F.

Stadler, editors. Foundations of Genetic Algorithms: 9th International Workshop, FOGA 2007, Mexico City, Mexico, January 8–11, 2007, Revised Selected Papers, volume 4436 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73479-1 (print), 3-540-73482-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73482-6.

Stojmenovic:2007:PDP

[STY^{+}07] Ivan Stojmenovic, Ruppa K. Thulasiram, Laurence T. Yang, Weijia Jia, Minyi Guo, and Rodrigo Fernandes de Mello, editors. Parallel and Distributed Processing and Applications: 5th International Symposium, ISPA 2007 Niagara Falls, Canada, August 29–31, 2007 Proceedings, volume 4742 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74741-9 (print), 3-540-74742-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74742-0.

Shyamasundar:1997:ACS

[SU97] Rudrapatna Shyamasundar and Kazunori Ueda, editors. Advances in computing science-ASIAN 97: Third Asian Computing Science Conference, Kathmandu, Nepal, December 1997: proceedings, volume 1345 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63875-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5.A735 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1345.
htm;   http://www.springerlink.
com/content/978-3-540-
63875-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1345.

Subsol:2005:VSU

[Sub05] Gérard Subsol, editor. Virtual Storytelling: Using Virtual Reality Technologies for Storytelling: Third International Conference, ICVS 2005, Strasbourg, France, November 30– December 2, 2005. Proceedings, volume 3805 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30511-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
30511-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3805.

Sudan:1995:ECP

[Sud95] Madhu Sudan. Efficient checking of polynomials and proofs and the hardness of approximation problems, volume 1001 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60615-7 (Berlin: softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 87 pp. LCCN QA267.S83 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1001.htm;
 http://www.springerlink.
com/content/978-3-540-
60615-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1001.

Suri:2002:MAI

[Sur02] Niranjan Suri, editor.

Mobile Agents: 6th International Conference, MA 2002, Barcelona, Spain, October 22–25, 2002. Proceedings, volume 2535 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00085-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 M3 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2535.htm;
http://www.springerlink.
com/content/978-3-540-
00085-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2535.

Schmeck:2002:TNP

[SUW02] Hartmut Schmeck, Theo Ungerer, and Lars Christian Wolf, editors. Trends in network and pervasive computing-ARCS 2002: International Conference on Architecture of Computing Systems, Karlsruhe, Germany, April 8–12, 2002: Proceedings, volume 2299 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43409-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A73 A736 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2299.htm;
 http://www.springerlink.
com/content/978-3-540-
43409-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2299.

Scholl:1997:ASD

[SV97] Michel O. Scholl and Agnes Voisard, editors. Advances in spatial databases: 5th international symposium, SSD ’97, Berlin, Germany, July 15–18, 1997: proceedings, volume 1262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63238-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.212.S69 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1262.
htm;   http://www.springerlink.
com/content/978-3-540-
63238-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1262.

Suciu:2001:WWW

[SV01] Dan Suciu and Gottfried Vossen, editors. The World Wide Web and databases: Third International Workshop WebDB 2000: Dallas, TX, USA, May 18– 19, 2000: Selected Papers, volume 1997 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41826-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.W43 W43 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1997.htm;
http://www.springerlink.
com/content/978-3-540-
41826-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1997.

Sutcliffe:2005:LPA

[SV05] Geoff Sutcliffe and Andrei Voronkov, editors. Logic for programming, artificial intelligence, and reasoning: 12th international conference, LPAR 2005, Montego Bay, Jamaica, December 2–6, 2005: proceedings, volume 3835 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30553-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .L73 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30553-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3835.

Szabo:2012:ICT

[SV12] Róbert Szabó and Attila Vidács, editors. Information and Communication Technologies: 18th EUNICE/IFIP WG 6.2, 6.6 International Conference, EUNICE 2012, Budapest, Hungary, August 29–31, 2012. Proceedings, volume 7479 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32807-5 (print), 3-642-32808-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32808-4.

Setubal:2005:ABC

[SVA05] João Carlos Setubal and Sergio Verjovski-Almeida, editors. Advances in Bioinformatics and Computational Biology: Brazilian Symposium on Bioinformatics, BSB 2005, São Leopoldo, Brazil, July 27–29, 2005: Proceedings, volume 3594 of Lecture Notes in Computer Science and Lecture Notes in Bioinformatics. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28008-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH441.2 .B73 2005. URL http:

//www.springerlink.
com/content/978-3-540-
28008-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3594.

Shi:2007:CSIa

[SvADS07a] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science — ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part I, volume 4487 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72583-0 (print), 3-540-72584-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72584-8.

Shi:2007:CSIb

[SvADS07b] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science — ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part II, volume 4488 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72585-7 (print), 3-540-72586-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72586-2.

Shi:2007:CSIc

[SvADS07c] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science — ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part III, volume 4489 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72587-3 (print), 3-540-72588-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72588-6.

Shi:2007:CSId

[SvADS07d] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science — ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part IV, volume 4490 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72589-X (print), 3-540-72590-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72590-9.

Sunderam:2005:CSIa

[SvASD05a] Vaidy S. Sunderam, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational science: ICCS 2005: 5th international conference, Atlanta, GA, USA, May 22–25, 2005, proceedings, Part I, volume 3514 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26032-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26032-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3514;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136570.

Sunderam:2005:CSIb

[SvASD05b] Vaidy S. Sunderam, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational science: ICCS 2005: 5th international conference, Atlanta, GA, USA, May 22–25, 2005, proceedings, Part II, volume 3515 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26043-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26043-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3515;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136571.

Sunderam:2005:CSIc

[SvASD05c] Vaidy S. Sunderam, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational science: ICCS 2005: 5th international conference, Atlanta, GA, USA, May 22–25, 2005, proceedings, Part III, volume 3516 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26044-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26044-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3516;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136575.

State:2006:LSM

[SvdMOP06] Radu State, Sven van der Meer, Declan O’Sullivan, and Tom Pfeifer, editors. Large Scale Management of Distributed Systems: 17th IFIP/IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2006, Dublin, Ireland, October 23–25, 2006. Proceedings, volume 4269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47659-8 (print), 3-540-47662-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        47662-7.

Scholten:2000:IDM

[SvS00] Hans Scholten and Martin J. van Sinderen, editors. Interactive Distributed Multimedia Systems and Telecommunication Services: 7th International Workshop, IDMS 2000, Enschede, The Netherlands, October 17–20, 2000. Proceedings, volume 1905 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41130-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1905. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1905.htm;
 http://www.springerlink.
com/content/978-3-540-
41130-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1905.

Sangiovanni-Vincentelli:2002:ESS

[SVS02]  Alberto  Sangiovanni-

Vincentelli and J. Sifakis, editors. Embedded Software: Second International Conference, EM-SOFT 2002, Grenoble, France, October 7– 9, 2002. Proceedings, volume 2491 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44307-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 E47 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2491.htm;
 http://www.springerlink.
com/content/978-3-540-
44307-0;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2491.

Sebillo:2008:VIS

[SVS08] Monica Sebillo, Giuliana Vitiello, and Gerald Schaefer, editors. Visual Information Systems. Web-Based Visual Information Search and Management: 10th International Conference, VISUAL 2008, Salerno, Italy, September 11– 12, 2008. Proceedings, volume 5188 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85890-3 (print), 3-540-85891-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85891-1.

Schurr:2012:AGT

[SVV12] Andreas Schürr, Dániel Varró, and Gergely Varró, editors. Applications of graph transformations with industrial relevance: 4th International Symposium, AG-TIVE 2011, Budapest, Hungary, October 4– 7, 2011, Revised selected and invited papers, volume 7233 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34175-6, 3-642-34176-4 (e-book). LCCN QA75.5 .A38 2011. URL http:

//www.springerlink.
com/content/978-3-642-
34175-5;   http://www.
springerlink.com/content/
978-3-642-34176-2.

Storer:1975:FAC

[SW75] Thomas Storer and David J. Winter, editors. Formal aspects of cognitive processes: [proceedings] interdisciplinary conference, Ann Arbor, March 1972, volume 22 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07016-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN BF311 .I561 1972. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0022.htm;
http://www.springerlink.
com/content/978-0-387-
07016-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=22.

Steinmetz:1997:IDM

[SW97] Ralf Steinmetz and Lars Christian Wolf, editors. Interactive distributed multimedia systems and telecommunication services: 4th International Workshop IDMS ’97, Darmstadt, Germany, September 1997:proceedings, volume 1309 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63519-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I59I586 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1309.
htm;   http://www.springerlink.
com/content/978-3-540-
63519-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1309.

Scholkopf:2003:LTK

[SW03a] Bernhard Schölkopf and Manfred K. Warmuth, editors. Learning Theory and Kernel Machines 16th Annual Conference on Learning Theory and 7th Kernel Workshop, COLT/ Kernel 2003, Washington, DC, USA, August 24–27, 2003: Proceedings, volume 2777 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40720-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2777.htm;
 http://www.springerlink.
com/content/978-3-540-
40720-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2777;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12006.

Schroeder:2003:RRM

[SW03b] Michael Schroeder and Gerd Wagner, editors. Rules and Rule Markup Languages for the Semantic Web: Second International Workshop, RuleML 2003, Sanibel Island, FL, USA, October 20, 2003: Proceedings, volume 2876 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20361-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.H94 R88 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2876.
htm;   http://www.springerlink.
com/content/978-3-540-
20361-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2876;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b94129.

Sahai:2004:UCI

[SW04] Akhil Sahai and Felix Wu, editors. Utility Computing: 15th IFIP/ IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2004, Davis, CA, USA, November 15–17, 2004. Proceedings, volume 3278 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23631-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3278.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23631-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3278;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b102082.

Steinmetz:2005:PPS

[SW05] Ralf Steinmetz and Klaus Wehrle, editors. Peer-to-peer systems and applications, volume 3485 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29192-X. ISSN 0302-9743 (print), 1611-3349 (electronic). xxvii + 629 pp. LCCN TK5105.525 .P442 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29192-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3485.

Sagot:2007:ABC

[SW07] Marie-France Sagot and Maria Emilia M. T. Walter, editors. Advances in Bioinformatics and Computational Biology: Second Brazilian Symposium on Bioinformatics, BSB 2007, Angra dos Reis, Brazil, August 29–31, 2007. Proceedings, volume 4643 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73730-8 (print), 3-540-73731-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73731-5.

Shankar:2008:VST

[SW08] Natarajan Shankar and Jim Woodcock, editors. Verified Software: Theories, Tools, Experiments: Second International Conference, VSTTE 2008, Toronto, Canada, October 6–9, 2008. Proceedings, volume 5295 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87872-6 (print), 3-540-87873-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87873-5.

Salzberg:2009:ABI

[SW09a] Steven L. Salzberg and Tandy Warnow, editors. Algorithms in Bioinformatics: 9th International Workshop, WABI 2009, Philadelphia, PA, USA, September 12–13, 2009. Proceedings, volume 5724 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04240-6 (print), 3-642-04241-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04241-6.

Shacham:2009:PBC

[SW09b] Hovav Shacham and Brent Waters, editors. Pairing-Based Cryptography — Pairing 2009: Third International Conference Palo Alto, CA, USA, August 12–14, 2009 Proceedings, volume 5671 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03297-4 (print), 3-642-03298-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03298-1.

Siegler:2010:VIT

[SW10] Simon Siegler and Nathan Wasser, editors. Verification, Induction, Termination Analysis: Festschrift for Christoph Walther on the Occasion of His 60th Birthday, volume 6463 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17171-0 (print), 3-642-17172-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17172-7.

Schubert:1995:HOL

[SWAF95] E. Thomas Schubert, Phillip J. Windley, and James Alves-Foss, editors. Higher order logic theorem proving and its applications: 8th International Workshop, Aspen Grove, UT, USA, September 11–14, 1995: proceedings, volume 971 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60275-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 H54 1995.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0971.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        60275-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=971.

Stevens:2003:FUU

[SWB03] Perdita Stevens, Jon Whittle, and Grady Booch, editors. ^{≪}UML^{≫}2003— The Unified Modeling Language: Modeling Languages and Applications 6th International Conference San Francisco, CA, USA, October 20–24, 2003 Proceedings, volume 2863 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20243-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2863.
htm;   http://www.springerlink.
com/content/978-3-540-
20243-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2863;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b14063.

Stanczak:2006:RAW

[SWB06] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche, editors. Resource Allocation in Wireless Networks: Theory and Algorithms, volume 4000 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-46248-1 (print), 3-540-46249-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

183 (est.) pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46249-1.

Swierstra:1999:PLS

[Swi99] S. Doaitse Swierstra, editor. Programming languages and systems: 8th European Symposium on Programming, ESOP’99, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS’99, Amsterdam, The Netherlands, March 22–28, 1999: Proceedings, volume 1576 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65699-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1576.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65699-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1576.

Sheng:2012:WTA

[SWJX12] Quan Z. Sheng, Guoren Wang, Christian S. Jensen, and Guandong Xu, editors. Web Technologies and Applications: 14th Asia-Pacific Web Conference, APWeb 2012, Kunming, China, April 11–13, 2012. Proceedings, volume 7235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29252-6 (print), 3-642-29253-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29253-8.

Stolen:2006:TMI

[SWMM06] Ketil Stølen, William H.

Winsborough, Fabio Martinelli, and Fabio Massacci, editors. Trust Management: 4th International Conference, iTrust 2006, Pisa, Italy, May 16–19, 2006. Proceedings, volume 3986 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34295-8 (print), 3-540-34297-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34297-7.

Slezak:2005:RSFa

[SWS^{+}05] Dominik Slezak, Guoyin Wang, Marcin Szczuka, Ivo Düntsch, and Yiyu Yao, editors. Rough Sets, Fuzzy Sets, Data Mining, and Granular Computing, volume 3641 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28653-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28653-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3641.

Suzuki:2011:MLM

[SWSY11] Kenji Suzuki, Fei Wang, Dinggang Shen, and Pingkun Yan, editors. Machine Learning in Medical Imaging: Second International Workshop, MLMI 2011, Held in Conjunction with MICCAI 2011, Toronto, Canada, September 18, 2011. Proceedings, volume 7009 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24318-5 (print), 3-642-24319-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24319-6.

Shi:2004:DMK

[SXC04] Yong Shi, Weixuan Xu, and Zhengxin Chen, editors. Data mining and knowledge management: Chinese Academy of Sciences Symposium CASDMKM 2004, Beijing, China, July 12– 14, 2004: Revised papers, volume 3327 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23987-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 C37 2004.

URL http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3327;
 http://www.springerlink.
com/content/978-3-540-
23987-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3327;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b104156.

Srikanthan:2005:ACS

[SXC05] Thambipillai Srikanthan, Jingling Xue, and Chip-Hong Chang, editors. Advances in Computer Systems Architecture: 10th Asia-Pacific Conference, ACSAC 2005, Singapore, October 24– 26, 2005. Proceedings, volume 3740 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29643-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29643-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3740.

Saitta:1988:UIS

[SYBM88] L. Saitta, Ronald R. Yager, and B. Bouchon-Meunier, editors. Uncertainty and intelligent systems, volume 313 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19402-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I5751 1988.

                                        URL http://link.springer-

ny.com/link/service/
series/0558/tocs/t0313.
htm;   http://www.springerlink.
com/content/978-0-387-
19402-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=313.

Samarati:2009:ISI

[SYMA09]  Pierangela  Samarati,

Moti Yung, Fabio Martinelli, and Claudio A. Ardagna, editors. Information Security: 12th International Conference, ISC 2009, Pisa, Italy, September 7–9, 2009. Proceedings, volume 5735 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04473-5 (print), 3-642-04474-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04474-8.

Salerno:2011:SCB

[SYNC11] John Salerno, Shanchieh Jay Yang, Dana Nau, and Sun-Ki Chai, editors. Social Computing, Behavioral-Cultural Modeling and Prediction: 4th International Conference, SBP 2011, College Park, MD, USA, March 29– 31, 2011. Proceedings, volume 6589 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19655-1 (print), 3-642-19656-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19656-0.

Slezak:2005:RSFb

[SYP^{+}05] Dominik Slezak, Jing-Tao Yao, James F. Peters, Wojciech Ziarko, and Xiaohua Hu, editors. Rough Sets, Fuzzy Sets, Data Mining, and Granular Computing: 10th International Conference, RSFDGrC 2005, Regina, Canada, August 31–September 3, 2005, Proceedings, Part II, volume 3642 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28660-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28660-8;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3642.

Syverson:2002:FCI

[Syv02] Paul F. Syverson, editor. Financial Cryptography: 5th International Conference, FC 2001, Grand Cayman, British West Indies, February 19–22, 2001. Proceedings, volume 2339 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44079-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2339.htm;
http://www.springerlink.
com/content/978-3-540-
44079-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2339.

Shen:2008:CSC

[SYY^{+}08] Weiming Shen, Jianming Yong, Yun Yang, Jean-Paul A. Barthès, and Junzhou Luo, editors. Computer Supported Cooperative Work in Design IV: 11th International Conference, CSCWD 2007, Melbourne, Australia, April 26–28, 2007. Revised Selected Papers, volume 5236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-92718-2 (print), 3-540-92719-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92719-8.

Segall:1992:DAI

[SZ92] Adrian Segall and Shmuel Zaks, editors. Distributed algorithms: 6th International Workshop, WDAG ’92, Haifa, Israel, November 2–4, 1992: proceedings, volume 647 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-56188-9 (Berlin), 0-387-56188-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D485 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0647.htm;
 http://www.springerlink.
com/content/978-3-540-
56188-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=647.

“The Sixth Workshop on Distributed Algorithms (WDAG 92)”–Pref.

Seberry:1993:ACA

[SZ93] Jennifer Seberry and Yuliang Zheng, editors. Advances in cryptology — AUSCRYPT ’92: Workshop on the Theory and Application of Cryptographic Techniques, Gold Coast, Queensland, Australia, December 13– 16, 1992: proceedings, volume 718 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-57220-1 (New York), 3-540-57220-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 A87 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0718.htm;
http://www.springerlink.
com/content/978-0-387-
57220-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=718.

Sommer:2000:AFP

[SZ00] Gerald Sommer and Y. Y. Zeevi, editors. Algebraic frames for the perception-action cycle: Second International Workshop, AF-PAC 2000, Kiel, Germany, September 10– 11, 2000: Proceedings, volume 1888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41013-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1888. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1888.htm;
 http://www.springerlink.
com/content/978-3-540-
41013-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1888.

Steffen:2005:FMO

[SZ05] Martin Steffen and Gianluigi Zavattaro, editors. Formal methods for open object-based distributed systems: 7th IFIP WG 6.1 international conference, FMOODS 2005, Athens, Greece, June 15–17, 2005: Proceedings, volume 3535 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26181-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26181-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3535;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136996.

Sturm:2011:ADG

[SZ11] Thomas Sturm and Christoph Zengler, editors. Automated Deduction in Geometry: 7th International Workshop, ADG 2008, Shanghai, China, September 22– 24, 2008. Revised Papers, volume 6301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21045-7 (print), 3-642-21046-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21046-4.

Sembok:2003:DLT

[SZC^{+}03] Tengku Mohd Tengku Sembok, Halimah Badioze Zaman, Hsinchun Chen, Shalini R. Urs, and Sung Hyon Myaeng, editors. Digital Libraries: Technology and Management of Indigenous Knowledge for Global Access: 6th International Conference on Asian Digital Libraries, ICADL 2003, Kuala Lumpur, Malaysia, December 8– 12, 2003 proceedings, volume 2911 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20608-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4080 .I578 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2911.htm;
 http://www.springerlink.
com/content/978-3-540-
20608-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2911;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94517.

Szepietowski:1994:TMS

[Sze94] Andrzej Szepietowski. Turing machines with sublogarithmic space, volume 843 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58355-6 (New York), 0-387-58355-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 114 pp. LCCN QA267 .S987 1994. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0843.
htm;   http://www.springerlink.
com/content/978-3-540-
58355-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=843.

Sandnes:2008:UIC

[SZR^{+}08] Frode Eika Sandnes, Yan Zhang, Chunming Rong, Laurence T. Yang, and Jianhua Ma, editors. Ubiquitous Intelligence and Computing: 5th International Conference, UIC 2008, Oslo, Norway, June 23–25, 2008 Proceedings, volume 5061 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69292-4 (print), 3-540-69293-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69293-5.

Spaccapietra:2009:JDSb

[SZS09] Stefano Spaccapietra, Esteban Zimányi, and Il-Yeol Song, editors. Journal on Data Semantics XIII, volume 5530 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03097-1 (print), 3-642-03098-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03098-7.

Sun:2008:ANNa

[SZT^{+}08a] Fuchun Sun, Jianwei Zhang, Ying Tan, Jinde Cao, and Wen Yu, editors. Advances in Neural Networks — ISNN 2008: 5th International Symposium on Neural Networks, ISNN 2008, Beijing, China, September 24–28, 2008, Proceedings, Part I, volume 5263 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87731-2 (print), 3-540-87732-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87732-5.

Sun:2008:ANNb

[SZT^{+}08b] Fuchun Sun, Jianwei Zhang, Ying Tan, Jinde Cao, and Wen Yu, editors. Advances in Neural Networks — ISNN 2008: 5th International Symposium on Neural Networks, ISNN 2008, Beijing, China, September 24–28, 2008, Proceedings, Part II, volume 5264 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87733-9 (print), 3-540-87734-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87734-9.

Taha:2000:SAI

[Tah00] Walid Taha, editor. Semantics, Applications, and Implementation of Program Generation: International Workshop, SAIG 2000, Montréal, Canada, September 20, 2000. Proceedings, volume 1924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41054-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S245 2000. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t1924.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41054-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1924.

Taha:2001:SAI

[Tah01] Walid Taha, editor. Semantics, applications, and implementation of program generation: Second International Workshop, SAIG 2001, Florence, Italy, September 6, 2001: Proceedings, volume 2196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42558-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S245 2001; QA267.A1 L43 no.2196. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2196.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42558-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2196.

Taha:2009:DSL

[Tah09] Walid Mohamed Taha, editor. Domain-Specific Languages: IFIP TC 2 Working Conference, DSL 2009 Oxford, UK, July 15–17, 2009 Proceedings, volume 5658 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03033-5 (print), 3-642-03034-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03034-5.

Tanguiane:1993:APM

[Tan93] Andranick S. Tanguiane.

Artificial perception and music recognition, volume 746 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57394-1 (Berlin), 0-387-57394-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 210 pp. LCCN ML74 .T36 1993. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0746.
htm;   http://www.springerlink.
com/content/978-3-540-
57394-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=746.

Tannier:2011:CGI

[Tan11] Eric Tannier, editor. Comparative Genomics: International Workshop, RECOMB-CG 2010, Ottawa, Canada, October 9–11, 2010. Proceedings, volume 6398 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-16180-4 (print), 3-642-16181-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16181-0.

Tarlecki:1991:MFC

[Tar91] Andrzej Tarlecki, editor. Mathematical foundations of computer science, 1991: 16th International Symposium, Kazimierz Dolny, Poland, September 9– 13, 1991: proceedings, volume 520 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54345-7 (Berlin), 0-387-54345-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.9.M35 M386 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0520.htm;
http://www.springerlink.
com/content/978-3-540-
54345-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=520.

Turker:2005:PPG

[TAS05] Can Türker, Maristella Agosti, and Hans-Jörg Schek, editors. Peer-to-peer, grid, and serviceorientation in digital library architectures: 6th thematic workshop of the EU network of excellence Delos, Cagliari, Italy, June 24–25, 2004. revised selected papers, volume 3664 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28711-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28711-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3664.

Taubner:1989:FRC

[Tau89] Dirk Taubner. Finite representations of CCS and TCSP programs by automata and Petri nets, volume 369 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-51525-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 168 pp. LCCN QA76.9.D5 T391 1989.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0369.
htm;   http://www.springerlink.
com/content/978-0-387-
51525-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=369.

Taubenfeld:2008:DCI

[Tau08] Gadi Taubenfeld, editor. Distributed Computing: 22nd International Symposium, DISC 2008, Arcachon, France, September 22–24, 2008. Proceedings, volume 5218 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87778-9 (print), 3-540-87779-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87779-0.

Thanos:2007:DLR

[TBC07] Costantino Thanos, Francesca Borri, and Leonardo Candela, editors. Digital Libraries: Research and Development: First International DELOS Conference, Pisa, Italy, February 13–14, 2007, Revised Selected Papers, volume 4877 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77087-9 (print), 3-540-77088-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77088-6.

Traff:2012:RAM

[TBD12] Jesper Larsson Träff, Siegfried Benkner, and Jack J. Dongarra, editors. Recent Advances in the Message Passing Interface: 19th European MPI Users’ Group Meeting, EuroMPI 2012, Vienna, Austria, September 23–26, 2012. Proceedings, volume 7490 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33517-9 (print), 3-642-33518-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33518-1.

Takizawa:2008:NBI

[TBE08] Makoto Takizawa, Leonard Barolli, and Tomoya Enokido, editors. Network-Based Information Systems: 2nd International Conference, NBiS 2008, Turin, Italy, September 1–5, 2008. Proceedings, volume 5186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85692-7 (print), 3-540-85693-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85693-1.

Tokuda:2009:PCI

[TBF^{+}09] Hideyuki Tokuda, Michael Beigl, Adrian Friday, A. J. Bernheim Brush, and Yoshito Tobe, editors. Pervasive Computing: 7th International Conference, Pervasive 2009, Nara, Japan, May 11–14, 2009. Proceedings, volume 5538 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01515-8 (print), 3-642-01516-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01516-8.

Tistarelli:2005:ASB

[TBG05] Massimo Tistarelli, Josef Bigun, and Enrico Grosso, editors. Advanced studies in biometrics: summer school on biometrics, Alghero, Italy, June 2–6, 2003. revised selected lectures and papers, volume 3161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26204-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26204-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3161;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136906.

Timmis:2003:AIS

[TBH03] Jon Timmis, Peter Bentley, and Emma Hart, editors. Artificial Immune Systems: Second International Conference, ICARIS 2003, Edinburgh, UK, September 1–3, 2003: Proceedings, volume 2787 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40766-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.875 .I23 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2787.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40766-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2787;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b12020.

Tistarelli:2002:BAI

[TBJ02] Massimo Tistarelli, Josef Bigun, and Anil K. Jain, editors. Biometric authentication: International ECCV 2002 Workshop, Copenhagen, Denmark, June 1, 2002: Proceedings, volume 2359 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43723-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 I568 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2359.htm;
 http://www.springerlink.
com/content/978-3-540-
43723-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2359.

Taylor:2010:SGI

[TBKO10] Robyn Taylor, Pierre Boulanger, Antonio Krüger, and Patrick Olivier, editors. Smart Graphics: 10th International Symposium on Smart Graphics, Banff, Canada, June 24–26, 2010 Proceedings, volume 6133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13543-9 (print), 3-642-13544-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13544-6.

Taylor:1995:SEH

[TC95] Richard N. Taylor and Joelle Coutaz, editors. Software engineering and human-computer interaction: ICSE ’94 Workshop on SE-HCI: Joint Research Issues, Sorrento, Italy, May 16– 17, 1994: proceedings, volume 896 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-59008-0 (Berlin), 0-387-59008-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I3 1994. DM62.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0896.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        59008-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=896.

Tombre:1998:GRA

[TC98] Karl Tombre and Atul K. Chhabra, editors. Graphics recognition: algorithms and systems: Second International Workshop, GREC 97, Nancy, France, August 1997: Selected Papers, volume 1389 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64381-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1637.I5538 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1389.htm;
 http://www.springerlink.
com/content/978-3-540-
64381-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1389.

Taylor:1999:MIC

[TC99] Chris (Christopher J.)

Taylor and A. C. F. (Alan C. F.) Colchester, editors. Medical image computing and computerassisted intervention — MICCAI ’99: Second International Conference, Cambridge, UK, September 19–22, 1999 proceedings, volume 1679 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66503-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN WB 141 I61m 1999; R857.O6 I574 1999. URL http:

//link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1679.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66503-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1679.

Tan:2012:AKDb

[TCHB12a] Pang-Ning Tan, Sanjay Chawla, Chin Kuan Ho, and James Bailey, editors. Advances in Knowledge Discovery and Data Mining: 16th Pacific-Asia Conference, PAKDD 2012, Kuala Lumpur, Malaysia, May 29 — June 1, 2012, Proceedings, Part II, volume 7302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30219-X (print), 3-642-30220-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30220-6.

Tan:2012:AKDa

[TCHB12b] Pang-Ning Tan, Sanjay Chawla, Chin Kuan Ho, and James Bailey, editors. Advances in Knowledge Discovery and Data Mining: 16th Pacific-Asia Conference, PAKDD 2012, Kuala Lumpur, Malaysia, May 29–June 1, 2012, Proceedings, Part I, volume 7301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30216-5 (print), 3-642-30217-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30217-6.

Touili:2010:CAV

[TCJ10] Tayssir Touili, Byron Cook, and Paul Jackson, editors. Computer Aided Verification: 22nd International Conference, CAV 2010, Edinburgh, UK, July 15–19, 2010. Proceedings, volume 6174 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14294-X (print), 3-642-14295-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14295-6.

Tsang:2010:HLT

[TCLH10] Philip Tsang, Simon K. S. Cheung, Victor S. K. Lee, and Ronghuai Huang, editors. Hybrid Learning: Third International Conference, ICHL 2010, Beijing, China, August 16–18, 2010. Proceedings, volume 6248 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14656-2 (print), 3-642-14657-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14657-2.

Theis:2012:LVA

[TCYZ12] Fabian Theis, Andrzej Cichocki, Arie Yeredor, and Michael Zibulevsky, editors. Latent Variable Analysis and Signal Separation: 10th International Conference, LVA/ ICA 2012, Tel Aviv, Israel, March 12–15, 2012. Proceedings, volume 7191 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28550-3 (print), 3-642-28551-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
28551-6.

tenCate:2007:LLC

[tCZ07] Balder D. ten Cate and Henk W. Zeevat, editors. Logic, Language, and Computation: 6th International Tbilisi Symposium on Logic, Language, and Computation,TbiLLC 2005 Batumi, Georgia, September 12–16, 2005. Revised Selected Papers, volume 4363 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75143-2 (print), 3-540-75144-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75144-1.

Taft:1997:ARM

[TD97] S. Tucker Taft and Robert A. Duff. Ada 95 reference manual: language and standard libraries, volume 1246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63144-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 526 pp. LCCN QA76.73.A35 A22 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1246.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63144-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1246.

International standard ISO/IEC 8652:1995(E).

Tesler:2007:CGR

[TD07] Glenn Tesler and Dannie Durand, editors. Comparative Genomics: RECOMB 2007 International Workshop, RECOMB-CG 2007, San Diego, CA, USA, September 16–18, 2007. Proceedings, volume 4751 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74959-4 (print), 3-540-74960-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74960-8.

Taft:2006:ARM

[TDB^{+}06] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, Erhard Ploedereder, and Pascal Leroy, editors. Ada 2005 Reference Manual. Language and Standard Libraries: International Standard ISO/ IEC 8652/1995 (E) with Technical Corrigendum 1 and Amendment 1, volume 4348 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69335-1 (print), 3-540-69336-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 618 (est.) pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69336-9.

Taft:2001:CAR

[TDBP01] S. Tucker Taft, R. A.

Duff, R. L. Brukardt, and E. Ploedereder, editors. Consolidated Ada reference manual: language and standard libraries: international standard ISO/IEC 8652/1995(E) with technical corrigendum 1, volume 2219 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-43038-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xix + 560 pp. LCCN QA76.73.A16 C65 2000. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2219.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43038-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2219.

Thalheim:1991:MSM

[TDG91] B. Thalheim, Janos Demetrovics, and H.- D. Gerhardt, editors. MFDBS 91: 3rd Symposium on Mathematical Fundamentals of Database and Knowledge Base Systems, Rostock, Germany, May 6–9, 1991: proceedings, volume 495 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 0-387-54009-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 S973 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0495.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        54009-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=495.

Trujillo:2010:ACM

[TDK^{+}10] Juan Trujillo, Gillian Dobbie, Hannu Kangassalo, Sven Hartmann, Markus Kirchberg, Matti Rossi, Iris Reinhartz-Berger, Esteban Zimányi, and Flavius Frasincar, editors. Advances in Conceptual Modeling — Applications and Challenges: ER 2010 Workshops ACM-L, CMLSA, CMS, DE@ER, FP-UML, SeCoGIS, WISM, Vancouver, BC, Canada, November 1–4, 2010. Proceedings, volume 6413 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16384-X (print), 3-642-16385-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16385-2.

Teixeira:2008:CPP

[TdLdOQ08] António Teixeira, Vera Lúcia Strube de Lima, Luís Caldas de Oliveira, and Paulo Quaresma, editors. Computational Processing of the Portuguese Language: 8th International Conference, PROPOR 2008 Aveiro, Portugal, September 8–10, 2008 Proceedings, volume 5190 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85979-9 (print), 3-540-85980-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85980-2.

Tscheligi:2009:AIE

[TdRM^{+}09] Manfred Tscheligi, Boris de Ruyter, Panos Markopoulus, Reiner Wichert, Thomas Mirlacher, Alexander Meschterjakov, and Wolfgang Reitberger, editors. Ambient Intelligence: European Conference, AmI 2009, Salzburg, Austria, November 18– 21, 2009. Proceedings, volume 5859 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05407-2 (print), 3-642-05408-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05408-2.

Tepfenhart:1994:CSC

[TDS94] William M. Tepfenhart, Judith P. Dick, and John F. Sowa, editors.

Conceptual structures, current practices: Second International Conference on Conceptual Structures, ICCS’94, College Park, Maryland, USA, August 16–20, 1994: proceedings, volume 835 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58328-9 (Berlin), 0-387-58328-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387 .I57 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0835.
htm;   http://www.springerlink.
com/content/978-3-540-
58328-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=835.

Takahashi:2011:EMC

[TDWG11] Ricardo H. C. Takahashi, Kalyanmoy Deb, Elizabeth F. Wanner, and Salvatore Greco, editors. Evolutionary Multi-Criterion Optimization: 6th International Conference, EMO 2011, Ouro Preto, Brazil, April 5– 8, 2011. Proceedings, volume 6576 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19892-9 (print), 3-642-19893-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19893-9.

Torsello:2009:GBR

[TEB09] Andrea Torsello, Francisco Escolano, and Luc Brun, editors. Graphbased representations in pattern recognition. 7th IAPR-TC-15 international workshop, GbRPR 2009, Venice, Italy, May 26–28, 2009. Proceedings, volume 5534 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. EUR 56.66.

Teillaud:1993:TDR

[Tei93] Monique Teillaud. Towards dynamic randomized algorithms in computational geometry, volume 758 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57503-0 (Berlin), 0-387-57503-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 156 pp. LCCN QA448.D38 T45 1993.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0758.
htm;   http://www.springerlink.
com/content/978-3-540-
57503-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=758.

terHaarRomeny:1997:SST

[ter97] Bart ter Haar Romeny, editor. Scale-space theory for computer vision: First International Conference, Scale-Space ’97, Utrecht, The Netherlands, July 2–4, 1997: proceedings, volume 1252 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63167-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .S38 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1252.htm;
 http://www.springerlink.
com/content/978-3-540-
63167-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1252.

Proceedings of the First International Conference on Scale-Space Theory in Computer Vision.

Tessaris:2009:RWS

[TFE^{+}09] Sergio Tessaris, Enrico Franconi, Thomas Eiter, Claudio Gutierrez, Siegfried Handschuh, Marie-Christine Rousset, and Renate A. Schmidt, editors. Reasoning Web. Semantic Technologies for Information Systems: 5th International Summer School 2009, Brixen-Bressanone, Italy, August 30 — September 4, 2009, Tutorial Lectures, volume 5689 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03753-4 (print), 3-642-03754-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03754-2.

Tanaka:2011:WWG

[TFK11] Katsumi Tanaka, Peter Fröhlich, and Kyoung-Sook Kim, editors. Web and Wireless Geographical Information Systems: 10th International Symposium, W2GIS 2011, Kyoto, Japan, March 3–4, 2011. Proceedings, volume 6574 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19172-X (print), 3-642-19173-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
19173-2.

Thomas:2000:HUC

[TG00] Peter J. Thomas and Hans-W W. Gellersen, editors. Handheld and Ubiquitous Computing: Second International Symposium, HUC 2000, Bristol, UK, September 25– 27, 2000. Proceedings, volume 1927 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41093-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1927. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1927.htm;
http://www.springerlink.
com/content/978-3-540-
41093-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1927.

Tomlin:2002:HSC

[TG02] Claire J. Tomlin and Mark Russell Greenstreet, editors. Hybrid systems: computation and control: 5th International Workshop, HSCC 2002, Stanford, CA, USA, March 25– 27, 2002: Proceedings, volume 2289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43321-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38 .H73 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2289.htm;
 http://www.springerlink.
com/content/978-3-540-
43321-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2289.

Tawfik:2004:AAI

[TG04] Ahmed Y. Tawfik and Scott D. Goodwin, editors. Advances in Artificial Intelligence: 17th Conference of the Canadian Society for Computational Studies of Intelligence, Canadian AI 2004, London, Ontario, Canada, May 17– 19, 2004: Proceedings, volume 3060 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22004-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .C36 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3060.htm;
 http://www.springerlink.
com/content/978-3-540-
22004-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3060;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97823.

Tratt:2010:TPM

[TG10] Laurence Tratt and Martin Gogolla, editors. Theory and Practice of Model Transformations: Third International Conference, ICMT 2010, Malaga, Spain, June 28– July 2, 2010. Proceedings, volume 6142 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13687-7 (print), 3-642-13688-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13688-7.

Timm:2012:MST

[TG12] Ingo J. Timm and Christian Guttmann, editors. Multiagent system technologies: 10th German Conference, MATES 2012, Trier, Germany, October 10–12, 2012. Proceedings, volume 7598 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33689-2, 3-642-33690-6 (e-book). LCCN QA76.76.I58 M38 2012.

                                        URL http://www.springerlink.
                                        com/content/978-3-642-
                                        33689-8;   http://www.
                                        springerlink.com/content/
                                        978-3-642-33690-4.

Travieso-Gonzalez:2011:ANS

[TGAH11] Carlos M. Travieso-González and Jesús B. Alonso-Hernández, editors. Advances in Nonlinear Speech Processing: 5th International Conference on Nonlinear Speech Processing, NO-LISP 2011, Las Palmas de Gran Canaria, Spain, November 7–9, 2011. Proceedings, volume 7015 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25019-X (print), 3-642-25020-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25020-0.

Thompson:2006:DAM

[TGH06] Simon G. Thompson and Robert Ghanea-Hercock, editors. Defence Applications of Multi-Agent Systems: International Workshop, DAMAS 2005, Utrecht, The Netherlands, July 25, 2005, Revised and Invited Papers, volume 3890 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32832-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32832-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3890.

Troccaz:1997:CMF

[TGM97] Jocelyne Troccaz, Eric Grimson, and Ralph Mosges, editors. CVRMed-MRCAS ’97: First Joint Conference Computer Vision, Virtual Reality and Robotics in Medicine and Medical Robotics and Computer-Assisted Surgery, Grenoble, France, March 19– 22, 1997: proceedings, volume 1205 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62734-0 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R859.7.C67 J65 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1205.htm;
 http://www.springerlink.
com/content/978-3-540-
62734-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1205.

Taniar:2010:CSAa

[TGM^{+}10a] David Taniar, Osvaldo Gervasi, Beniamino Murgante, Eric Pardede, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2010: International Conference, Fukuoka, Japan, March 23–26, 2010, Proceedings, Part I, volume 6016 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12155-1 (print), 3-642-12156-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12156-2.

Taniar:2010:CSAb

[TGM^{+}10b] David Taniar, Osvaldo Gervasi, Beniamino Murgante, Eric Pardede, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2010: International Conference, Fukuoka, Japan, March 23–26, 2010, Proceedings, Part II, volume 6017 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12164-0 (print), 3-642-12165-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12165-4.

Taniar:2010:CSAc

[TGM^{+}10c] David Taniar, Osvaldo Gervasi, Beniamino Murgante, Eric Pardede, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2010: International Conference, Fukuoka, Japan, March 23–26, 2010, Proceedings, Part III, volume 6018 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12178-0 (print), 3-642-12179-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12179-1.

Taniar:2010:CSAd

[TGM^{+}10d] David Taniar, Osvaldo Gervasi, Beniamino Murgante, Eric Pardede, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2010: International Conference, Fukuoka, Japan, March 23–26, 2010, Proceedings, Part IV, volume 6019 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12188-8 (print), 3-642-12189-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12189-0.

Thayse:1981:BCD

[Tha81] Andre Thayse. Boolean calculus of differences, volume 101 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10286-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 143 pp. LCCN QA10.3 .T47. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0101.htm;
 http://www.springerlink.
com/content/978-0-387-
10286-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=101.

Thayse:1984:PFB

[Tha84]  Andre Thayse.  P-

functions and Boolean matrix factorization: a unified approach for wired, programmed, and microprogrammed implementations of discrete algorithms, volume 175 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-13358-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 248 pp. LCCN QA76.9.S88 T44. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0175.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        13358-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=175.

Thalheim:1996:CMI

[Tha96] B. (Bernhard) Thalheim, editor. Conceptual modeling — ER ’96: 15th International Conference on Conceptual Modeling, Cottbus, Germany, October 7–10, 1996: proceedings, volume 1157 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61784-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D26 I567 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1157.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61784-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1157.

terHofstede:2008:BPM

[tHBP08] Arthur ter Hofstede, Boualem Benatallah, and Hye-Young Paik, editors. Business Process Management Workshops: BPM 2007 International Workshops, BPI, BPD, CBP, Pro-Health, RefMod, semantics4ws, Brisbane, Australia, September 24, 2007, Revised Selected Papers, volume 4928 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78237-0 (print), 3-540-78238-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78238-4.

Theoharis:1989:APP

[The89] T. (Theoharis) Theoharis. Algorithms for parallel polygon rendering, volume 373 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 3-540-51394-9, 0-387-51394-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 147 pp. LCCN QA76.5 .T521 1989. DM33.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0373.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        51394-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=373.

Thiagarajan:1994:FST

[Thi94] P. S. Thiagarajan, editor. Foundations of software technology and theoretical computer science: 14th conference, Madras, India, December 15–17, 1994: proceedings, volume 880 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58715-2 (New York), 3-540-58715-2 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F692 1994. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0880.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        58715-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=880.

Thiagarajan:1995:FST

[Thi95] P. S. Thiagarajan, editor. Foundations of software technology and theoretical computer science: 15th conference, Bangalore, India, December 18–20, 1995: proceedings, volume 1026 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60692-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F692 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1026.htm;
 http://www.springerlink.
com/content/978-3-540-
60692-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1026.

Thielscher:2000:CAT

[Thi00a] Michael Thielscher. Challenges for action theories, volume 1775 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67455-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 138 pp. LCCN QA76.9.L63 T48 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1775.htm;
 http://www.springerlink.
com/content/978-3-540-
67455-9;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1775.

Thierauf:2000:CCE

[Thi00b] Thomas Thierauf. The Computational Complexity of Equivalence and Isomorphism Problems, volume 1852 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41032-5. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 135 pp. LCCN QA169 .T45 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1852.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41032-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1852.

Thilikos:2010:GTC

[Thi10] Dimitrios M. Thilikos, editor. Graph Theoretic Concepts in Computer Science: 36th International Workshop, WG 2010, Zarós, Crete, Greece, June 28–30, 2010 Revised Papers, volume 6410 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16925-2 (print), 3-642-16926-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16926-7.

Tang:2011:IUK

[THL11] Yongchuan Tang, Van-Nam Huynh, and Jonathan Lawry, editors. Integrated Uncertainty in Knowledge Modelling and Decision Making: International Symposium, IUKM 2011, Hangzhou, China, October 28–30, 2011. Proceedings, volume 7027 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24917-5 (print), 3-642-24918-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24918-1.

Thomas:1999:FSS

[Tho99] Wolfgang Thomas, editor. Foundations of software science and computation structures: Second International Conference, FOSSACS ’99 held as part of the joint European Conferences on Theory and Practice of Software, ETAPS ’99, Amsterdam, The Netherlands, March 22– 28, 1999: Proceedings, volume 1578 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65719-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F67 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1578.htm;
 http://www.springerlink.
com/content/978-3-540-
65719-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1578.

Thomas:2006:EOO

[Tho06] Dave Thomas, editor. ECOOP 2006 — Object-Oriented Programming: 20th European Conference, Nantes, France, July 3–7, 2006. Proceedings, volume 4067 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35726-2 (print), 3-540-35727-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35727-8.

Thomas:2011:CPE

[Tho11] Nigel Thomas, editor. Computer Performance Engineering: 8th European Performance Engineering Workshop, EPEW 2011, Borrowdale, UK, October 12– 13, 2011. Proceedings, volume 6977 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24748-2 (print), 3-642-24749-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24749-1.

Truszkowski:2003:ICA

[THR03] Walt Truszkowski, Mike Hinchey, and Chris Rouff, editors. Innovative Concepts for Agent-Based Systems: First International Workshop on Radical Agent Concepts, WRAC 2002, McLean, VA, USA, January 16– 18, 2002: Revised Papers, volume 2564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40725-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 I594 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2564.
htm;   http://www.springerlink.
com/content/978-3-540-
40725-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2564;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b83557.

Tyrrell:2003:ESB

[THT03] Andy M. Tyrrell, Pauline C. Haddow, and Jim Torresen, editors. Evolvable systems: from biology to hardware: 5th International Conference, ICES 2003, Trondheim, Norway, March 17–20, 2003: Proceedings, volume 2606 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00730-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .I57 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2606.htm;
 http://www.springerlink.
com/content/978-3-540-

00730-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2606.

Thulasiraman:2007:FHP

[THX^{+}07] Parimala Thulasiraman, Xubin He, Tony Li Xu, Mieso K. Denko, Ruppa K. Thulasiram, and Laurence T. Yang, editors. Frontiers of High Performance Computing and Networking ISPA 2007 Workshops: ISPA 2007 International Workshops SS-DSN, UPWN, WISH, SGC, ParDMCom, HiP-CoMB, and IST-AWSN Niagara Falls, Canada, August 28–September 1, 2007 Proceedings, volume 4743 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74766-4 (print), 3-540-74767-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74767-3.

Tison:1994:TAP

[Tis94] Sophie Tison, editor.

Trees in algebra and programming, CAAP’94: 19th international colloquium, Edinburgh, UK, April 11–13, 1994: proceedings, volume 787 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57879-X (Berlin), 0-387-57879-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166.2 .I58 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0787.htm;
 http://www.springerlink.
com/content/978-3-540-
57879-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=787.

19th Colloquium on Trees in Algebra and Programming (CAAP ’94).

Tison:2002:RTA

[Tis02] Sophie Tison, editor. Rewriting Techniques and Applications: 13th International Conference, RTA 2002, Copenhagen, Denmark, July 22–24, 2002. Proceedings, volume 2378 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43916-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I555 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2378.htm;
 http://www.springerlink.
com/content/978-3-540-
43916-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2378.

Tiuryn:2000:FSS

[Tiu00] Jerzy Tiuryn, editor.

Foundations of software science and computation structures: Third International Conference, FOSSACS 2000, held as part of the Joint European Conference on Theory and Practice of Software, ETAPS 2000, Berlin, Germany, March 25–April 2, 2000: Proceedings, volume 1784 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67257-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1784. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1784.htm;
 http://www.springerlink.
com/content/978-3-540-
67257-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1784.

Thomas:2008:CPE

[TJ08] Nigel Thomas and Carlos Juiz, editors. Computer Performance Engineering: 5th European Performance Engineering Workshop, EPEW 2008, Palma de Mallorca, Spain, September 24–25, 2008. Proceedings, volume 5261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87411-9 (print), 3-540-87412-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87412-6.

Takasu:1994:LLC

[TJHS94] Satoru Takasu, Neil D. Jones, Masami Hagiya, and Masahiko Sato, editors. Logic, language, and computation: Festschrift in honor of Satoru Takasu, volume 792 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-57935-4 (New York), 3-540-57935-4 (Berlin).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .L56625 1994. DM50.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0792.
htm;   http://www.springerlink.
com/content/978-0-387-
57935-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=792.

Tergan:2005:KIV

[TK05] Sigmar-Olaf Tergan and Tanja Keller, editors. Knowledge and Information Visualization: Searching for Synergies, volume 3426 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26921-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.I6 K66 2005.

URL http://www.springerlink.
com/content/978-3-540-
26921-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3426.

Theeramunkong:2009:AKD

[TKCH09] Thanaruk Theeramunkong, Boonserm Kijsirikul, Nick Cercone, and Tu-Bao Ho, editors. Advances in Knowledge Discovery and Data Mining: 13th Pacific-Asia Conference, PAKDD 2009 Bangkok, Thailand, April 27–30, 2009 Proceedings, volume 5476 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01306-6 (print), 3-642-01307-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01307-2.

Tang:2011:ADMa

[TKCW11a] Jie Tang, Irwin King, Ling Chen, and Jianyong Wang, editors. Advanced Data Mining and Applications: 7th International Conference, ADMA 2011, Beijing, China, December 17–19, 2011, Proceedings, Part I, volume 7120 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25852-2 (print), 3-642-25853-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25853-4.

Tang:2011:ADMb

[TKCW11b] Jie Tang, Irwin King, Ling Chen, and Jianyong Wang, editors. Advanced Data Mining and Applications: 7th International Conference, ADMA 2011, Beijing, China, December 17–19, 2011, Proceedings, Part II, volume 7121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25855-7 (print), 3-642-25856-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25856-5.

Treharne:2005:ZFS

[TKHS05] Helen Treharne, Steve King, Martin Henson, and Steve Schneider, editors. ZB 2005: formal specification and development in Z and B: 4th International Conference of B and Z Users, Guildford, UK, April 13–15, 2005: Proceedings, volume 3455 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25559-1 (paperback).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25559-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3455;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b135596.

Theeramunkong:2011:KIC

[TKSN11] Thanaruk Theeramunkong, Susumu Kunifuji, Virach Sornlertlamvanich, and Cholwich Nattee, editors. Knowledge, Information, and Creativity Support Systems: 5th International Conference, KICSS 2010, Chiang Mai, Thailand, November 25–27, 2010, Revised Selected Papers, volume 6746 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24787-3 (print), 3-642-24788-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24788-0.

Thalheim:1998:SD

[TL98] B. (Bernhard) Thalheim and Leonid Libkin. Semantics in databases, volume 1358 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64199-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 265 pp. LCCN QA76.9.D3S438 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1358.htm;
 http://www.springerlink.
com/content/978-3-540-
64199-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1358.

Traunmuller:2002:EGF

[TL02] Roland Traunmüller and Klaus Lenk, editors. Electronic government: First International Conference, EGOV 2002, Aix-en-Provence, France, September 2–6, 2002: Proceedings, volume 2456 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44121-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN JF1525.A8 E584 2002. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t2456.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44121-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2456.

Terano:2000:KDD

[TLC00] Takao Terano, Huan Liu, and Arbee L. P. Chen, editors. Knowledge discovery and data mining: current issues and new applications: 4th Pacific-Asia Conference, PAKDD 2000, Kyoto, Japan, April 18–20, 2000: Proceedings, volume 1805 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67382-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1805.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1805.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67382-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1805.

Talbi:2006:AEI

[TLC06] El-Ghazali Talbi, Pierre Liardet, and Pierre Collet, editors. Artificial evolution: 7th international conference, evolution artificielle, EA 2005, Lille, France, October 26–28, 2005, revised selected papers, volume 3871 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33590-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33590-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3871.

Tan:2001:MDM

[TLF01] Kian-Lee Tan, John Chi-Shing Lui, and Michael J. Franklin, editors. Mobile data management: Second International Conference, MDM 2001, Hong Kong, China, January 2001: Proceedings, volume 1987 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41454-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.59 .I58 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1987.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41454-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1987.

Todorovski:2006:DSI

[TLJ06] Ljupčo Todorovski, Nada Lavrač, and Klaus P. Jantke, editors. Discovery Science: 9th International Conference, DS 2006, Barcelona, Spain, October 7–10, 2006. Proceedings, volume 4265 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46491-3 (print), 3-540-46493-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        46493-8.

Tang:2008:ADM

[TLZ^{+}08] Changjie Tang, Charles X. Ling, Xiaofang Zhou, Nick J. Cercone, and Xue Li, editors. Advanced Data Mining and Applications: 4th International Conference, ADMA 2008, Chengdu, China, October 8–10, 2008. Proceedings, volume 5139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88191-3 (print), 3-540-88192-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88192-6.

Tavares:1999:SAC

[TM99] Stafford Tavares and Henk Meijer, editors. Selected areas in cryptography: 5th annual International Workshop, SAC ’98, Kingston, Ontario, Canada, August 17–18, 1998: Proceedings, volume 1556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65894-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 S22 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1556.htm;
 http://www.springerlink.
com/content/978-3-540-
65894-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1556.

Takagi:2009:AIC

[TM09] Tsuyoshi Takagi and Masahiro Mambo, editors. Advances in Information and Computer Security: 4th International Workshop on Security, IWSEC 2009 Toyama, Japan, October 28–30, 2009 Proceedings, volume 5824 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04845-5 (print), 3-642-04846-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        04846-3.

Tambouris:2011:EPT

[TMdB11] Efthimios Tambouris, Ann Macintosh, and Hans de Bruijn, editors. Electronic Participation: Third IFIP WG 8.5 International Conference, ePart 2011, Delft, The Netherlands, August 29 — September 1, 2011. Proceedings, volume 6847 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23332-5 (print), 3-642-23333-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23333-3.

Tambouris:2010:EPS

[TMG10] Efthimios Tambouris, Ann Macintosh, and Olivier Glassey, editors. Electronic Participation: Second IFIP WG 8.5 International Conference, ePart 2010, Lausanne, Switzerland, August 29 — September 2, 2010. Proceedings, volume 6229 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15157-4 (print), 3-642-15158-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15158-3.

Tai:2009:SSV

[TMLL09] Xue-Cheng Tai, Knut Mørken, Marius Lysaker, and Knut-Andreas Lie, editors. Scale Space and Variational Methods in Computer Vision: Second International Conference, SSVM 2009, Voss, Norway, June 1–5, 2009. Proceedings, volume 5567 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02255-3 (print), 3-642-02256-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02256-2.

Trigoni:2009:GNT

[TMN09] Niki Trigoni, Andrew Markham, and Sarfraz Nawaz, editors. GeoSensor Networks: Third International Conference, GSN 2009, Oxford, UK, July 13–14, 2009. Proceedings, volume 5659 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02902-7 (print), 3-642-02903-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02903-5.

Trinder:2004:IFL

[TMP04] Phil Trinder, Greg Michaelson, and Ricardo Peäna, editors. Implementation of functional languages:

15th international workshop, IFL 2003, Edinburgh, UK, September 8– 11, 2003: revised papers, volume 3145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23727-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.62 .I56 2003. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3145;
 http://www.springerlink.
com/content/978-3-540-
23727-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3145;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b102274.

Tambouris:2012:EPI

[TMS12] Efthimios Tambouris, Ann Macintosh, and Øystein Sæbø, editors. Electronic Participation: 4th IFIP WG 8.5 International Conference, ePart 2012, Kristiansand, Norway, September 3– 5, 2012. Proceedings, volume 7444 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33249-8 (print), 3-642-33250-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33250-0.

Taylor:2003:IPM

[TN03] Chris (Christopher J.) Taylor and J. Alison Noble, editors. Information Processing in Medical Imaging: 18th International Conference, IPMI 2003 Ambleside, UK, July 20–25, 2003 Proceedings, volume 2732 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40560-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN RC78.7.D53 I573 2003.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2732.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        40560-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2732;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11820.

Torra:2004:MDA

[TN04] Vicenç Torra and Yasuo Narukawa, editors. Modeling Decisions for Artificial Intelligence: First International Conference, MDAI 2004, Barcelona, Catalonia, Spain, August 2–4, 2004: Proceedings, volume 3131 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22555-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q335. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3131.htm;
http://www.springerlink.
com/content/978-3-540-
22555-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3131;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99254.

Torra:2008:MDA

[TN08] Vicenç Torra and Yasuo Narukawa, editors. Modeling Decisions for Artificial Intelligence: 5th International Conference, MDAI 2008 Sabadell, Spain, October 30–31, 2008. Proceedings, volume 5285 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88268-5 (print), 3-540-88269-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88269-5.

Tistarelli:2009:ABT

[TN09] Massimo Tistarelli and Mark S. Nixon, editors. Advances in Biometrics: Third International Conference, ICB 2009, Alghero, Italy, June 2–5, 2009. Proceedings, volume 5558 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01792-4 (print), 3-642-01793-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01793-3.

Theeramunkong:2010:NFA

[TNA^{+}10] Thanaruk Theeramunkong, Cholwich Nattee, Paulo J. L. Adeodato, Nitesh Chawla, Peter Christen, Philippe Lenca, Josiah Poon, and Graham Williams, editors. New Frontiers in Applied Data Mining: PAKDD 2009 International Workshops, Bangkok, Thailand, April 27–30, 2009. Revised Selected Papers, volume 5669 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14639-2 (print), 3-642-14640-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14640-4.

Torra:2010:MDA

[TND10] Vicenç Torra, Yasuo Narukawa, and Marc Daumas, editors. Modeling Decisions for Artificial Intelligence: 7th International Conference, MDAI 2010, Perpignan, France, October 27–29, 2010. Proceedings, volume 6408 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16291-6 (print), 3-642-16292-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16292-3.

Tuyls:2008:AAM

[TNGK08] Karl Tuyls, Ann Nowe, Zahia Guessoum, and Daniel Kudenko, editors. Adaptive Agents and Multi-Agent Systems III. Adaptation and Multi-Agent Learning: 5th, 6th, and 7th European Symposium, ALAMAS 2005-2007 on Adaptive and Learning Agents and Multi-Agent Systems, Revised Selected Papers, volume 4865 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77947-7 (print), 3-540-77949-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77949-0.

Torra:2009:MDA

[TNI09] Vicenç Torra, Yasuo Narukawa, and Masahiro Inuiguchi, editors. Modeling Decisions for Artificial Intelligence: 6th International Conference, MDAI 2009, Awaji Island, Japan, November 30–December 2, 2009. Proceedings, volume 5861 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04819-6 (print), 3-642-04820-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04820-3.

Torra:2005:MDA

[TNM05] Vicenç Torra, Yasuo Narukawa, and Sadaaki Miyamoto, editors. Modeling Decisions for Artificial Intelligence: Second International Conference, MDAI 2005, Tsukuba, Japan, July 25–27, 2005. Proceedings, volume 3558 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27871-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .M33 2005. URL http:

//www.springerlink.
com/content/978-3-540-
27871-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3558.

Terano:2001:NFA

[TNN^{+}01] Takao Terano, T. Nishida, A. Namatame, S. Tsumoto, Y. Ohsawa, and T. Washio, editors. New Frontiers in Artificial Intelligence: Joint JSAI 2001 Workshop Post-Proceedings, volume 2253 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43070-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .N49 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2253.htm;
 http://www.springerlink.
com/content/978-3-540-
43070-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2253.

Tomkos:2007:OND

[TNP^{+}07] Ioannis Tomkos, Fabio Neri, Josep Solé Pareta, Xavier Masip Bruin, and Sergi Sánchez Lopez, editors. Optical Network Design and Modeling: 11th International IFIP TC6 Conference, ONDM 2007, Athens, Greece, May 29–31, 2007. Proceedings, volume 4534 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72729-9 (print), 3-540-72731-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72731-6.

Torra:2006:MDA

[TNVDF06] Vicenç Torra, Yasuo Narukawa, Aïda Valls, and Josep Domingo-Ferrer, editors. Modeling Decisions for Artificial Intelligence: Third International Conference, MDAI 2006, Tarragona, Spain, April 3–5, 2006. Proceedings, volume 3885 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32780-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32780-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3885.

Torra:2007:MDA

[TNY07] Vicenç Torra, Yasuo Narukawa, and Yuji Yoshida, editors. Modeling Decisions for Artificial Intelligence: 4th International Conference, MDAI 2007, Kitakyushu, Japan, August 16–18, 2007. Proceedings, volume 4617 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73728-6 (print), 3-540-73729-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73729-2.

Torra:2011:MDA

[TNYL11] Vicenç Torra, Yasuo Narakawa, Jianping Yin, and Jun Long, editors. Modeling Decision for Artificial Intelligence: 8th International Conference, MDAI 2011, Changsha, Hunan, China, July 28–30, 2011, Proceedings, volume 6820 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22588-8 (print), 3-642-22589-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22589-5.

Tokunaga:2008:LSK

[TO08] Takenobu Tokunaga and Antonio Ortega, editors. Large-Scale Knowledge Resources. Construction and Application: Third International Conference on Large-Scale Knowledge Resources, LKR 2008, Tokyo, Japan, March 3–5, 2008. Proceedings, volume 4938 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78158-7 (print), 3-540-78159-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78159-2.

Tobias:1980:LDP

[Tob80] Jeffrey M. Tobias, editor.

Language design and programming methodology: proceedings of a symposium held at Sydney, Australia, 10–11 September 1979, volume 79 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CODEN LNCSD9. ISBN 0-387-09745-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .S94 1979. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0079.htm;
 http://www.springerlink.
com/content/978-0-387-

                                        09745-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=79.

Sponsored by the Australian Atomic Energy Commission and the University of New South Wales.

Tokuyama:2007:ACI

[Tok07] Takeshi Tokuyama, editor. Algorithms and Computation: 18th International Symposium, ISAAC 2007, Sendai, Japan, December 17– 19, 2007. Proceedings, volume 4835 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77118-2 (print), 3-540-77120-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77120-3.

Tscheligi:2008:CTE

[TOL08] Manfred Tscheligi, Marianna Obrist, and Artur Lugmayr, editors. Changing Television Environments: 6th European Conference, EU-ROITV 2008, Salzburg, Austria, July 3–4, 2008 Proceedings, volume 5066 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69477-3 (print), 3-540-69478-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69478-6.

Tomayko:1991:SEE

[Tom91] J. E. Tomayko, editor.

Software engineering education: SEI Conference 1991, Pittsburgh, Pennsylvania, USA, October 7–8, 1991, proceedings, volume 536 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 0-387-54502-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S45 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0536.htm;
 http://www.springerlink.
com/content/978-0-387-
54502-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=536.

“The Fifth SEI Conference on Software Engineering”–Pref.

Tomek:1992:CAL

[Tom92] Ivan Tomek, editor. Computer assisted learning: 4th International Conference, ICCAL ’92, Wolfville, Nova Scotia, Canada, June 17– 20, 1992: proceedings, volume 602 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55578-1 (Berlin), 0-387-55578-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1028.5 .I11635 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0602.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55578-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=602.

Takagi:2007:PBC

[TOOO07] Tsuyoshi Takagi, Tatsuaki Okamoto, Eiji Okamoto, and Takeshi Okamoto, editors. Pairing-Based Cryptography — Pairing 2007: First International Conference, Tokyo, Japan, July 2– 4, 2007. Proceedings, volume 4575 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73488-0 (print), 3-540-73489-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
73489-5.

Torasso:1993:AAI

[Tor93] Pietro Torasso, editor. Advances in artificial intelligence: Third Congress of the Italian Association for Artificial Intelligence, AI*IA ’93, Torino, Italy, October 26–28, 1993: proceedings, volume 728 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57292-9 (New York), 3-540-57292-9 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I83 1993. DM66.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0728.htm;
 http://www.springerlink.
com/content/978-0-387-
57292-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=728.

Toussaint:1994:AEF

[Tou94] M. Toussaint, editor. Ada in Europe: First International Eurospace-Ada-Europe Symposium, Copenhagen, Denmark, September 26–30, 1994: proceedings, volume 887 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58822-1 (Berlin: paperback), 0-387-58822-1 (New York: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 I57 1994. DM60.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0887.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58822-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=887.

Toussaint:1996:AES

[Tou96] M. Toussaint, editor. Ada in Europe: Second International Eurospace — Ada-Europe Symposium, Frankfurt/ Main, Germany, October 2–6, 1995: proceedings, volume 1031 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60757-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 I57 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1031.htm;
 http://www.springerlink.
com/content/978-3-540-
60757-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1031.

Tokoro:1994:OOP

[TP94] Mario Tokoro and Remo Pareschi, editors. Objectoriented programming: 8th European Conference, ECOOP ’94, Bologna, Italy, July 4–8, 1994: proceedings, volume 821 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-58202-9 (New York), 3-540-58202-9 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1994. DM96.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0821.
htm;   http://www.springerlink.
com/content/978-0-387-
58202-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=821.

Trappl:1997:CPS

[TP97] Robert Trappl and Paolo Petta, editors. Creating personalities for synthetic actors: towards autonomous personality agents, volume 1195 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62735-9 (Berlin softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TR897.7 .C7 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1195.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62735-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1195.

Tollis:2009:GDI

[TP09] Ioannis G. Tollis and Maurizio Patrignani, editors. Graph Drawing: 16th International Symposium, GD 2008, Heraklion, Crete, Greece, September 21–24, 2008. Revised Papers, volume 5417 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00218-8 (print), 3-642-00219-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00219-9.

Takasu:1979:MSI

[TPB79] Satoru Takasu, M. Paul, and E. K. Blum, editors. Mathematical studies of information processing: proceedings of the international conference, Kyoto, Japan, August 23– 26, 1978, volume 75 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CO-DEN LNCSD9. ISBN 0-387-09541-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I582 1978. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0075.htm;
http://www.springerlink.
com/content/978-0-387-
09541-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=75.

Thilagam:2012:ACN

[TPCB12] P. Santhi Thilagam, Alwyn Roshan Pais, K. Chandrasekaran, and N. Balakrishnan, editors. Advanced Computing, Networking and Security: International Conference, ADCONS 2011, Surathkal, India, December 16–18, 2011, Revised Selected Papers, volume 7135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29279-8 (print), 3-642-29280-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29280-4.

Tjoa:2011:ARS

[TQYX11] A. Min Tjoa, Gerald Quirchmayr, Ilsun You, and Lida Xu, editors. Availability, Reliability and Security for Business, Enterprise and Health Information Systems: IFIP WG 8.4/ 8.9 International Cross Domain Conference and Workshop, ARES 2011, Vienna, Austria, August 22–26, 2011. Proceedings, volume 6908 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23299-X (print), 3-642-23300-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23300-5.

Taft:2012:PAM

[TR12] Nina Taft and Fabio Ricciato, editors. Passive and Active Measurement: 13th International Conference, PAM 2012, Vienna, Austria, March 12–14th, 2012. Proceedings, volume 7192 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28536-8 (print), 3-642-28537-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28537-0.

Traunmuller:2004:EGS

[Tra04a] Roland Traunmüller, editor. Electronic Government: Second International Conference, EGOV 2003, Prague, Czech Republic, September 1–5, 2003: Proceedings, volume 2739 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-40845-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2739.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40845-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2739;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11827.

Traunmuller:2004:EGT

[Tra04b] Roland Traunmüller, editor. Electronic Government: Third International Conference, EGOV 2004, Zaragoza, Spain, August 30–September 3, 2004. Proceedings, volume 3183 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22916-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3183.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22916-2;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3183;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99836.

Treinen:2009:RTA

[Tre09] Ralf Treinen, editor.

Rewriting Techniques and Applications: 20th International Conference, RTA 2009 Brasília, Brazil, June 29 — July 1, 2009 Proceedings, volume 5595 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02347-9 (print), 3-642-02348-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02348-4.

Triglia:1998:ISN

[Tri98] Sebastiano Triglia, editor. Intelligence in services and networks: technology for ubiquitous telecom services: 5th International Conference on Intelligence in Services and Networks, IS and N ’98, Antwerp, Belgium, May 25–28, 1998: Proceedings, volume 1430 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64598-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1430. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1430.htm;
 http://www.springerlink.
com/content/978-3-540-
64598-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1430.

Troubitsyna:2011:SER

[Tro11] Elena A. Troubitsyna, editor. Software Engineering for Resilient Systems: Third International Workshop, SERENE 2011, Geneva, Switzerland, September 29–30, 2011. Proceedings, volume 6968 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24123-9 (print), 3-642-24124-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24124-6.

Tinhofer:1987:GTC

[TS87] Gottfried Tinhofer and Gunther Schmidt, editors. Graph-theoretic concepts in computer science: International Workshop WG ’86, Bernried, Federal Republic of Germany, June 17–19, 1986: proceedings, volume 246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-17218-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G7161 1987. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0246.htm;
http://www.springerlink.
com/content/978-0-387-
17218-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=246.

Thai:2010:CCA

[TS10] My T. Thai and Sartaj Sahni, editors. Computing and Combinatorics: 16th Annual International Conference, COCOON 2010, Nha Trang, Vietnam, July 19–21, 2010. Proceedings, volume 6196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14030-0 (print), 3-642-14031-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14031-0.

Tan:2011:ASIa

[TSCW11a] Ying Tan, Yuhui Shi, Yi Chai, and Guoyin Wang, editors. Advances in Swarm Intelligence: Second International Conference, ICSI 2011, Chongqing, China, June 12–15, 2011, Proceedings, Part I, volume 6728 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21514-9 (print), 3-642-21515-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21515-5.

Tan:2011:ASIb

[TSCW11b] Ying Tan, Yuhui Shi, Yi Chai, and Guoyin Wang, editors. Advances in Swarm Intelligence: Second International Conference, ICSI 2011, Chongqing, China, June 12–15, 2011, Proceedings, Part II, volume 6729 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21523-8 (print), 3-642-21524-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21524-7.

Tomkos:2009:TDO

[TSE^{+}09] Ioannis Tomkos, Maria Spyropoulou, Karin Ennser, Martin Köhn, and Branko Mikac, editors. Towards Digital Optical Networks: COST Action 291 Final Report, volume 5412 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01523-9 (print), 3-642-01524-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01524-3.

Tan:2000:AMI

[TSG00] Tieniu Tan, Yuanchun Shi, and Wen Gao, editors. Advances in Multimodal Interfaces—ICMI 2000: Third International Conference, Beijing, China, October 14–16, 2000. Proceedings, volume 1948 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41180-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.U83 I26 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1948.htm;
 http://www.springerlink.
com/content/978-3-540-
41180-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1948.

Tan:2012:ASIa

[TSJ12a] Ying Tan, Yuhui Shi, and Zhen Ji, editors. Advances in Swarm Intelligence: Third International Conference, ICSI 2012, Shenzhen, China, June 17–20, 2012 Proceedings, Part I, volume 7331 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30975-5 (print), 3-642-30976-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
30976-2.

Tan:2012:ASIb

[TSJ12b] Ying Tan, Yuhui Shi, and Zhen Ji, editors. Advances in Swarm Intelligence: Third International Conference, ICSI 2012, Shenzhen, China, June 17–20, 2012 Proceedings, Part II, volume 7332 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31019-2 (print), 3-642-31020-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31020-1.

Toueg:1992:DAI

[TSK92] S. Toueg, P. G. Spirakis, and L. Kirousis, editors. Distributed algorithms: 5th International Workshop, WDAG ’91, Delphi, Greece, October 7–9, 1991: proceedings, volume 579 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55236-7, 0-387-55236-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D4853 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0579.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55236-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=579.

“Fifth International Workshop on Distributed Algorithms … ”–Pref.

Tsumoto:2004:RSC

[TSKGB04] Shusaku Tsumoto, Roman Słowiński, Jan Komorowski, and Jerzy W. Grzymała-Busse, editors. Rough Sets and Current Trends in Computing: 4th International Conference, RSCTC 2004, Uppsala, Sweden, June 1–5, 2004: Proceedings, volume 3066 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22117-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .R. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3066.htm;
                                         http://www.springerlink.

com/content/978-3-540-
22117-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3066;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97961.

Tinelli:2011:FCS

[TSS11] Cesare Tinelli and Viorica Sofronie-Stokkermans, editors. Frontiers of Combining Systems: 8th International Symposium, FroCoS 2011, Saarbrücken, Germany, October 5–7, 2011. Proceedings, volume 6989 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24363-0 (print), 3-642-24364-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24364-6.

Tremeau:2009:CCI

[TST09] Alain Trémeau, Raimondo Schettini, and Shoji Tominaga, editors. Computational Color Imaging: Second International Workshop, CCIW 2009, Saint-Etienne, France, March 26–27, 2009. Revised Selected Papers, volume 5646 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03264-8 (print), 3-642-03265-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03265-3.

Tan:2010:ASIa

[TST10a] Ying Tan, Yuhui Shi, and Kay Chen Tan, editors. Advances in Swarm Intelligence: First International Conference, ICSI 2010, Beijing, China, June 12–15, 2010, Proceedings, Part I, volume 6145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13494-7 (print), 3-642-13495-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13495-1.

Tan:2010:ASIb

[TST10b] Ying Tan, Yuhui Shi, and Kay Chen Tan, editors. Advances in Swarm Intelligence: First International Conference, ICSI 2010, Beijing, China, June 12–15, 2010, Proceedings, Part II, volume 6146 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13497-1 (print), 3-642-13498-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13498-2.

Tsudik:2008:FCD

[Tsu08] Gene Tsudik, editor.

Financial Cryptography and Data Security: 12th International Conference, FC 2008, Cozumel, Mexico, January 28–31, 2008. Revised Selected Papers, volume 5143 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85229-8 (print), 3-540-85230-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85230-8.

Tamassia:1995:GDD

[TT95] Roberto Tamassia and Ioannis G. Tollis, editors. Graph drawing: DIMACS International Workshop, GD ’94, Princeton, New Jersey, USA, October 10–12, 1994: proceedings, volume 894 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-58950-3 (Berlin), 0-387-58950-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .D55 1994. DM94.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0894.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58950-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=894.

Tjoa:2005:DWK

[TT05] A. Min Tjoa and Juan Trujillo, editors. Data warehousing and knowledge discovery: 7th international conference, DaWaK 2005, Copenhagen, Denmark, August 22–26, 2005: proceedings, volume 3589 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28558-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28558-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3589.

Tjoa:2006:DWK

[TT06a] A. Min Tjoa and Juan Trujillo, editors. Data Warehousing and Knowledge Discovery: 8th International Conference, DaWaK 2006, Krakow, Poland, September 4– 8, 2006. Proceedings, volume 4081 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37736-0 (print), 3-540-37737-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37737-5.

Toni:2006:CLM

[TT06b] Francesca Toni and Paolo Torroni, editors. Computational Logic in Multi-Agent Systems: 6th International Workshop, CLIMA VI, London, UK, June 27– 29, 2005: Revised Selected and Invited Papers, volume 3900 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33996-5 (paperback). LCCN QA76.9.L63 C67 2005. URL http:

//www.springerlink.
com/content/978-3-540-
33996-0.

Taylor:2010:ALA

[TT10] Matthew E. Taylor and Karl Tuyls, editors. Adaptive and Learning Agents: Second Workshop, ALA 2009, Held as Part of the AAMAS 2009 Conference in Budapest, Hungary, May 12, 2009. Revised Selected Papers, volume 5924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11813-5 (print), 3-642-11814-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11814-2.

Tovar:2007:PDS

[TTF07] Eduardo Tovar, Philippas Tsigas, and Hacène Fouchal, editors. Principles of Distributed Systems: 11th International Conference, OPODIS 2007, Guadeloupe, French West Indies, December 17–20, 2007. Proceedings, volume 4878 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77095-X (print), 3-540-77096-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77096-1.

Tuyls:2006:LAM

[TtHVS06] Karl Tuyls, Pieter Jan ’t Hoen, Katja Verbeeck, and Sandip Sen, editors. Learning and Adaption in Multi-Agent Systems: First International Workshop, LAMAS 2005, Utrecht, The Netherlands, July 25, 2005, Revised Selected Papers, volume 3898 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33053-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33053-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3898.

Tempesti:2010:ESB

[TTM10] Gianluca Tempesti, Andy M. Tyrrell, and Julian F. Miller, editors. Evolvable Systems: From Biology to Hardware: 9th International Conference, ICES 2010, York, UK, September 6–8, 2010. Proceedings, volume 6274 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15322-4 (print), 3-642-15323-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15323-5.

Tao:2005:ACI

[TTP05] Jianhua Tao, Tieniu Tan, and Rosalind W. Picard, editors. Affective computing and intelligent interaction: first international conference, ACII 2005, Beijing, China, October 22–24, 2005: proceedings, volume 3784 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29621-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 A25 2005.

URL http://www.springerlink.
com/content/978-3-540-
29621-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3784.

tenTeije:2012:KEK

[tTVH^{+}12] Annette ten Teije, Johanna Völker, Siegfried Handschuh, Heiner Stuckenschmidt, Mathieu d’Acquin, Andriy Nikolov, Nathalie Aussenac-Gilles, and Nathalie Hernandez, editors. Knowledge Engineering and Knowledge Management: 18th International Conference, EKAW 2012, Galway City, Ireland, October 8–12, 2012. Proceedings, volume 7603 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33875-5 (print), 3-642-33876-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        33876-2.

Tomas:1994:VSP

[TU94] Gerald Tomas and Christoph W. Ueberhuber. Visualization of scientific parallel programs, volume 771 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-57738-6 (Berlin), 0-387-57738-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 310 pp. LCCN QA76.642 .V57 1994.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0771.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        57738-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=771.

Tu:1999:AAC

[Tu99] Xiaoyuan Tu. Artificial Animals for Computer Animation: Biomechanics, Locomotion, Perception, and Behavior, volume 1635 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66939-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 172 pp. LCCN TR897.7. T8 1999. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1635.
htm;   http://www.springerlink.
com/content/978-3-540-
66939-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1635.

Originally presented as the author’s thesis (doctoral)– University of Toronto, 1996.

Tucci:2001:MDI

[Tuc01] Maurizio Tucci, editor.

Multimedia databases and image communication: Second International Workshop, MDIC 2001, Amalfi, Italy, September 17–18, 2001: Proceedings, volume 2184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42587-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .M36 2001; QA267.A1 L43 no.2184. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2184.htm;
 http://www.springerlink.
com/content/978-3-540-
42587-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=2184.

Treleaven:1987:FPC

[TV87] P. C. Treleaven and Marco Vanneschi, editors. Future parallel computers: an advanced course, Pisa, Italy, June 9–10, 1986: proceedings, volume 272 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 0-387-18203-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .F881 1987. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0272.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        18203-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=272.

Proceedings of Advanced Course on Future Parallel Computers. “Organized by the University of Pisa on behalf of the European Strategic Programme for Research and Development in Information Technology (ESPRIT), sponsored by Philips, The Netherlands … [et al.]”– Verso of t.p.

Tel:1994:DAI

[TV94] Gerard Tel and Paul M. B. Vitanyi, editors. Distributed algorithms: 8th International Workshop / WDAG ’94, Terschelling, the Netherlands, September 29– October 1, 1994, proceedings, volume 857 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58449-8 (Berlin), 0-387-58449-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I597 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0857.htm;
http://www.springerlink.
com/content/978-3-540-
58449-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=857.

Tanabe:2002:DCI

[TvdBI02] Makoto Tanabe, Peter van den Besselaar, and Toru Ishida, editors. Digital cities II: Computational and sociological approaches: second Kyoto Workshop on Digital Cities, Kyoto, Japan, October 18– 20, 2001: Revised Papers, volume 2362 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43963-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.83 .K96 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2362.htm;
 http://www.springerlink.
com/content/978-3-540-
43963-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2362.

Thomas:2007:SAS

[TW07] Wolfgang Thomas and Pascal Weil, editors. STACS 2007: 24th Annual Symposium on Theoretical Aspects of Computer Science, Aachen, Germany, February 22– 24, 2007. Proceedings, volume 4393 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70917-7 (print), 3-540-70918-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70918-3.

Thilikos:2012:PEC

[TW12] Dimitrios M. Thilikos and Gerhard J. Woeginger, editors. Parameterized and Exact Computation: 7th International Symposium, IPEC 2012, Ljubljana, Slovenia, September 12– 14, 2012. Proceedings, volume 7535 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33292-7 (print), 3-642-33293-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33293-7.

Tokoro:1992:OBC

[TWN92] Mario Tokoro, Peter Wegner, and Oscar Marius Nierstrasz, editors. Object-based concurrent computing: ECOOP ’91 workshop, Geneva, Switzerland, July 15– 16, 1991: proceedings, volume 612 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55613-3 (Berlin), 0-387-55613-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0612.htm;
 http://www.springerlink.
com/content/978-3-540-
55613-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=612.

Selected presentations from the ECOOP’91 Workshop on Object-Based Concurrent Computing held in Geneva, Switzerland, July 15–16, 1991.

Takagi:2012:PSI

[TWQ^{+}12] Tsuyoshi Takagi, Guilin Wang, Zhiguang Qin, Shaoquan Jiang, and Yong Yu, editors. Provable Security: 6th International Conference, ProvSec 2012, Chengdu, China, September 26– 28, 2012. Proceedings, volume 7496 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33271-4 (print), 3-642-33272-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33272-2.

Tuyls:2007:KDE

[TWSN07] Karl Tuyls, Ronald Westra, Yvan Saeys, and Ann Nowé, editors. Knowledge Discovery and Emergent Complexity in Bioinformatics: First International Workshop, KDECB 2006, Ghent, Belgium, May 10, 2006. Revised Selected Papers, volume 4366 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71036-1 (print), 3-540-71037-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71037-0.

Tzschach:1977:TCS

[TWW77] H. G. Tzschach, Helmut Waldschmidt, and H. K.-G. Walter, editors. Theoretical computer science, volume 48 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CO-DEN LNCSD9. ISBN 0-387-08138-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.G181 1977. URL

http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t0048.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        08138-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=48.

English, French, or Ger-

man.

Tang:2001:WAA

[TWYhL01] Yuan Y. Tang, Victor Wickerhauser, Pong C. Yuen, and Chun hung Li, editors. Wavelet Analysis and Its Applications: Second International Conference, WAA 2001, Hong Kong, China, December 18– 20, 2001: Proceedings, volume 2251 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43034-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA403.3 .W32 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2251.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43034-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2251.

Taylor:2011:IPC

[TY11] Russell H. Taylor and Guang-Zhong Yang, editors. Information Processing in Computer-Assisted Interventions: Second International Conference, IPCAI 2011, Berlin, Germany, June 22, 2011. Proceedings, volume 6689 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21503-3 (print), 3-642-21504-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21504-9.

Tsumoto:2005:AMS

[TYNM05] Shusaku Tsumoto, Takahira Yamaguchi, Masayuki Numao, and Hiroshi Motoda, editors. Active mining: second international workshop, AM 2003, Maebashi, Japan, October 28, 2003. Revised selected papers, volume 3430 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26157-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        26157-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3430;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136140.

Temam:2011:APP

[TYZ11] Olivier Temam, Pen-Chung Yew, and Binyu Zang, editors. Advanced Parallel Processing Technologies: 9th International Symposium, APPT 2011, Shanghai, China, September 26– 27, 2011. Proceedings, volume 6965 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24150-6 (print), 3-642-24151-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24151-2.

Torn:1989:GO

[TZ89] Aimo Torn and A. Zhilinskas. Global optimization, volume 350 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 3-540-50871-6 (Berlin), 0-387-50871-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 255 pp. LCCN QA402.5 .T631 1989.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0350.
htm;   http://www.springerlink.
com/content/978-3-540-
50871-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=350.

Triggs:2000:VAT

[TZS00] Bill Triggs, Andrew Zisserman, and Richard Szeliski, editors. Vision Algorithms: Theory and Practice: International Workshop on Vision Algorithms, Corfu, Greece, September 21–22, 1999. Proceedings, volume 1883 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67973-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .I72 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1883.htm;
 http://www.springerlink.
com/content/978-3-540-
67973-8;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1883.

USDOD:1983:PLA

[UA83] United States. Dept. of Defense and American National Standards Institute. The Programming language Ada reference manual, volume 155 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-12328-8, 3-540-12328-8. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 330 pp. LCCN QA76.73.A35 P72 1983. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0155.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        12328-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=155.

“ANSI/MIL-STD-1815A-1983.” Ada was designed in accordance with requirements defined by the United States Dept. of Defense. “Approved 17 February 1983.”.

Unger:2002:IIC

[UBM02] Herwig Unger, Thomas Böhme, and Armin Mikler, editors. Innovative Internet Computing Systems: Second International Workshop, IICS 2002, Kühlungsborn, Germany, June 20–22, 2002. Proceedings, volume 2346 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43790-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I52 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2346.htm;
 http://www.springerlink.
com/content/978-3-540-
43790-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2346.

Ur:2006:HSV

[UBW06] Shmuel Ur, Eyal Bin, and Yaron Wolfsthal, editors. Hardware and Software, Verification and Testing: First International Haifa Verification Conference, Haifa, Israel, November 13–16, 2005, Revised Selected Papers, volume 3875 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32604-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        32604-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3875.

Unay:2010:RPS

[ÜÇA10] Devrim Ünay, Zehra Çataltepe, and Selim Aksoy, editors. Recognizing Patterns in Signals, Speech, Images and Videos: ICPR 2010 Contests, Istanbul, Turkey, August 23– 26, 2010, Contest Reports, volume 6388 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17710-7 (print), 3-642-17711-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17711-8.

Uyar:2006:TCS

[UDF06] M. Uemit Uyar, Ali Y. Duale, and Mariusz A. Fecko, editors. Testing of Communicating Systems: 18th IFIP TC 6/WG 6.1 International Conference, TESTCOM 2006, New York, NY, USA, May 16–18, 2006, Proceedings, volume 3964 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34184-6 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34184-0.

Ueda:2010:PLS

[Ued10] Kazunori Ueda, editor. Programming Languages and Systems: 8th Asian Symposium, APLAS 2010, Shanghai, China, November 28 — December 1, 2010. Proceedings, volume 6461 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17163-X (print), 3-642-17164-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17164-2.

Uhl:1982:AGS

[Uhl82] J. Uhl. An Attribute grammar for the semantic analysis of ADA, volume 139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11571-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 511 pp. LCCN QA76.73.A35 .A87 1982. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0139.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        11571-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=139.

Ullrich:2008:PFC

[Ull08] Carsten Ullrich, editor. Pedagogically Founded Courseware Generation for Web-Based Learning: An HTN-Planning-Based Approach Implemented in PAIGOS, volume 5260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CO-DEN LNCSD9. ISBN 3-540-88213-8 (print), 3-540-88215-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 232 (est.) pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        88215-2.

Umeo:2008:CAI

[UMN^{+}08] Hiroshi Umeo, Shin Morishita, Katsuhiro Nishinari, Toshihiko Komatsuzaki, and Stefania Bandini, editors. Cellular Automata: 8th International Conference on Cellular Aotomata for Reseach and Industry, ACRI 2008, Yokohama, Japan, September 23–26, 2008. Proceedings, volume 5191 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79991-5 (print), 3-540-79992-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79992-4.

USDOD:1981:PLA

[Uni81] United States. Dept. of Defense. The Programming Language Ada Reference Manual: Proposed Standard Document, volume 106 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10693-6, 3-540-10693-6. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 243 pp. LCCN QA76.73.A35 P763 1981. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0106.htm;
 http://www.springerlink.

                                        com/content/978-0-387-
                                        10693-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=106.

A photographic reproduction of the official November 1980 printing (Honeywell, Minneapolis).

Uhlig:2007:PAN

[UPB07] Steve Uhlig, Konstantina Papagiannaki, and Olivier Bonaventure, editors. Passive and Active Network Measurement: 8th International Conference, PAM 2007, Louvainla-neuve, Belgium, April 5–6, 2007. Proceedings, volume 4427 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71616-5 (print), 3-540-71617-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        71617-4.

Ursino:2002:EEI

[Urs02] Domenico Ursino. Extraction and Exploitation of Intensional Knowledge from Heterogeneous Information Sources: Semi-Automatic Approaches and Tools, volume 2282 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43347-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xxvi + 289 pp. LCCN QA76.88 .U77 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2282.htm;
 http://www.springerlink.
com/content/978-3-540-
43347-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2282.

Originally presented as the author’s thesis (Universita degli Studi della Calabria).

Urzyczyn:2005:TLC

[Urz05] Paweł Urzyczyn, editor. Typed lambda calculi and applications: 7th international conference, TLCA 2005, Nara, Japan, April 21– 23, 2005: Proceedings, volume 3461 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25593-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.5 .I58. URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        25593-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3461;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b135690.

Uhl:1990:SCR

[US90] J. (Jurgen) Uhl and Hans Albrecht Schmid. A systematic catalogue of reusable abstract data types, volume 460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-53229-3, 0-387-53229-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 344 pp. LCCN QA76.9.A23 U37 1990. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0460.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        53229-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=460.

Uustalu:2006:MPC

[Uus06] Tarmo Uustalu, editor. Mathematics of Program Construction: 8th International Conference, MPC 2006, Kuressaare, Estonia, July 3–5, 2006.

Proceedings, volume 4014 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35631-2 (print), 3-540-35632-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35632-5.

Umeda:2006:DPK

[UWB^{+}06] Masanobu Umeda, Armin Wolf, Oskar Bartenstein, Ulrich Geske, Dietmar Seipel, and Osamu Takata, editors. Declarative Programming for Knowledge Management: 16th International Conference on Applications of Declarative Programming and Knowledge Management, INAP 2005, Fukuoka, Japan, October 22–24, 2005. Revised Selected Papers, volume 4369 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-69233-9 (print), 3-540-69234-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        69234-8.

Vieira:2006:CPP

[V^{+}06] Renata Vieira et al., editors. Computational Processing of the Portuguese Language: 7th International Workshop, PROPOR 2006, Itatiaia, Brazil, May 13–17, 2006: Proceedings, volume 3960 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34045-9 (paperback). LCCN PC5044.5 .I58 2006. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34045-4.

Veloso:1995:CBR

[VA95] Manuela Veloso and Agnar Aamodt, editors. Case-based reasoning research and development: First International Conference, ICCBR-95, Sesimbra, Portugal, October 23–26, 1995: proceedings, volume 1010 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60598-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 C376 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1010.htm;
http://www.springerlink.
com/content/978-3-540-
60598-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1010.

Veit:2007:GEB

[VA07] Daniel J. Veit and Jörn Altmann, editors. Grid Economics and Business Models: 4th International Workshop, GECON 2007, Rennes, France, August 28, 2007. Proceedings, volume 4685 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74428-2 (print), 3-540-74430-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74430-6.

Vaananen:1999:GQC

[Vää99] J. Väänänen, editor.

Generalized quantifiers and computation: 9th European Summer School in Logic, Language, and Information, ESS-LLI’97 workshop: Aixen-Provence, France, August 11–22, 1997: revised lectures, volume 1754 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66993-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 E97 1997.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1754.
htm;   http://www.springerlink.
com/content/978-3-540-
66993-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1754.

Vadhan:2007:TCT

[Vad07] Salil P. Vadhan, editor. Theory of Cryptography: 4th Theory of Cryptography Conference, TCC 2007, Amsterdam, The Netherlands, February 21–24, 2007. Proceedings, volume 4392 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70935-5 (print), 3-540-70936-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
70936-7.

Vahrenhold:2009:EAI

[Vah09] Jan Vahrenhold, editor. Experimental Algorithms: 8th International Symposium, SEA 2009, Dortmund, Germany, June 4–6, 2009. Proceedings, volume 5526 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02010-0 (print), 3-642-02011-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02011-7.

Valette:1994:ATP

[Val94] Robert Valette, editor.

Application and theory of Petri nets, 1994: 15th International Conference, Zaragoza, Spain, June 20–24, 1994: proceedings, volume 815 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58152-9 (Berlin), 0-387-58152-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I56 1994.

URL http://link.springer-

                                        ny.com/link/service/
                                        series/0558/tocs/t0815.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        58152-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=815.

“Proceedings of the 15th International Conference on Application and Theory of Petri Nets”–Pref.

Valmari:2006:MCS

[Val06] Antti Valmari, editor. Model Checking Software: 13th International SPIN Workshop, Vienna, Austria, March 30–April 1, 2006. Proceedings, volume 3925 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-33102-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33102-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3925.

Vounckx:2006:ICS

[VAM06] Johan Vounckx, Nadine Azemard, and Philippe Maurine, editors. Integrated Circuit and System Design. Power and Timing Modeling, Optimization and Simulation: 16th International Workshop, PAT-MOS 2006, Montpellier, France, September 13– 15, 2006. Proceedings, volume 4148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39094-4 (print), 3-540-39097-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
39097-8.

VandeSnepscheut:1985:TTV

[Van85] Jan L. A. Van de Snepscheut. Trace theory and VLSI design, volume 200 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1985. CO-DEN LNCSD9. ISBN 0-387-15988-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 140 pp. LCCN TK7874 .S6251 1985. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0200.htm;
 http://www.springerlink.
com/content/978-0-387-
15988-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=200.

Thesis (Ph. D.) — Eindhoven University of Technology.

VandeSnepscheut:1989:MPC

[Van89] Jan L. A. Van de Snepscheut, editor. Mathematics of program construction: 375th anniversary of the Groningen University: international conference, Groningen, The Netherlands, June 26–30, 1989: proceedings, volume 375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51305-1 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .M3661 1989. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0375.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51305-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=375.

Papers presented at the Conference on Mathematics of Program Construction.

vanGasteren:1990:SMA

[van90] A. J. M. (Antonetta J. M.) van Gasteren. On the shape of mathematical arguments, volume 445 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 3-540-52849-0 (Berlin), 0-387-52849-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 180 pp. LCCN QA9.54 .G37 1990. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0445.
htm;   http://www.springerlink.
com/content/978-3-540-
52849-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=445.

VanLoocke:1994:DCC

[Van94] Philip R. Van Loocke.

The dynamics of concepts: a connectionist model, volume 766 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57647-9 (Berlin: softcover), 0-387-57647-9 (New York: softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 340 pp. LCCN Q325.5 .L66 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0766.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        57647-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=766.

VanHentenryck:1997:SAI

[Van97a] Pascal Van Hentenryck, editor. Static analysis: 4th International Symposium, SAS’97, Paris, France, September 1997: proceedings, volume 1302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63468-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6.I5733 1997. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1302.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63468-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1302.

vanKreveld:1997:AFG

[van97b] Marc van Kreveld. Algorithmic foundations of geographic information systems, volume 1340 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63818-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 287 pp. LCCN G70.212.A38 1997. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1340.
htm;   http://www.springerlink.
com/content/978-3-540-
63818-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1340.

VanHentenryck:2002:PPC

[Van02] Pascal Van Hentenryck, editor. Principles and Practice of Constraint Programming — CP 2002: 8th International Conference, CP 2002, Ithaca, NY, USA, September 9–13, 2002: Proceedings, volume 2470 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44120-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C7 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2470.htm;
 http://www.springerlink.
com/content/978-3-540-
44120-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2470.

VanRoy:2005:MPM

[Van05] Peter Van Roy, editor. Multiparadigm programming in Mozart/ OZ: second international conference, MOZ 2004, Charleroi, Belgium, October 7–8, 2004. Revised selected and invited papers, volume 3389 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25079-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 . M95 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25079-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3389;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106627.

Vanmechelen:2012:EGC

[VAR12] Kurt Vanmechelen, Jörn Altmann, and Omer F. Rana, editors. Economics of Grids, Clouds, Systems, and Services: 8th International Workshop, GECON 2011, Paphos, Cyprus, December 5, 2011, Revised Selected Papers, volume 7150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28674-7 (print), 3-642-28675-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28675-9.

Vouros:2009:OAM

[VASP09] George Vouros, Alexander Artikis, Kostas Stathis, and Jeremy Pitt, editors. Organized Adaption in Multi-Agent Systems: First International Workshop, OA-MAS 2008, Estoril, Portugal, May 13, 2008. Revised and Invited Papers, volume 5368 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02376-2 (print), 3-642-02377-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02377-4.

Vaudenay:1998:FSE

[Vau98] Serge Vaudenay, editor.

Fast software encryption: 5th International Workshop, FSE ’98, Paris, France, March 23–25, 1998: Proceedings, volume 1372 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64265-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25F77 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1372.htm;
 http://www.springerlink.
com/content/978-3-540-
64265-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1372.

Vaudenay:2005:PKC

[Vau05] Serge Vaudenay, editor. Public key cryptography — PKC 2005: 8th International Workshop on Theory and Practice in Public Key Cryptography, Les Diablerets, Switzerland, January 23–26, 2005: Proceedings, volume 3386 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24454-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I567 2005. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3386;   http:
//www.springerlink.
com/content/978-3-540-
24454-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3386;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105124.

Vaudenay:2006:ACE

[Vau06] Serge Vaudenay, editor. Advances in Cryptology — EUROCRYPT 2006: 25th International Conference on the Theory and Applications of Cryptographic Techniques, St. Petersburg, Russia, May 28– June 1, 2006, Proceedings, volume 4004 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34547-7. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34547-3.

Vaudenay:2008:PCA

[Vau08] Serge Vaudenay, editor.

Progress in Cryptology — AFRICACRYPT 2008: First International Conference on Cryptology in Africa, Casablanca, Morocco, June 11–14, 2008. Proceedings, volume 5023 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68159-0 (print), 3-540-68164-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68164-9.

Vazirgiannis:1999:IMD

[Vaz99] Michalis Vazirgiannis. Interactive Multimedia Documents: Modeling, Authoring, and Implementation Experiences, volume 1564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66711-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 161 pp. LCCN QA76.76.I59 V39 1999; QA267.A1 L43 no.1564. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1564.htm;
 http://www.springerlink.
com/content/978-3-540-
66711-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1564.

vanDalen:1997:CSL

[vB97] D. van Dalen and M. Bezem, editors. Computer science logic: 10th International Workshop, CSL’96, Annual Conference of the EACSL, Utrecht, The Netherlands, September 21– 27, 1996: Selected Papers, volume 1258 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63172-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.L63 W67 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1258.htm;
http://www.springerlink.
com/content/978-3-540-
63172-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1258.

vanBeek:2005:PPC

[vB05] Peter G. van Beek, editor. Principles and practice of constraint programming – CP 2005: 11th international conference, CP 2005, Sitges, Spain, October 1– 5, 2005: proceedings, volume 3709 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29238-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C7 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29238-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3709.

vanBreugel:2008:CCT

[vBC08] Franck van Breugel and Marsha Chechik, editors. CONCUR 2008

— Concurrency Theory: 19th International Conference, CONCUR 2008, Toronto, Canada, August 19–22, 2008. Proceedings, volume 5201 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85360-X (print), 3-540-85361-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85361-9.

Vojtas:2005:STP

[VBCBS05] Peter Vojtáš, Mária Bieliková, Bernadette Charron-Bost, and Ondrej Sýkora, editors. SOFSEM 2005: theory and practice of computer science: 31st conference on current trends in theory and practice of computer science, Liptovský Ján, Slovakia, January 22–28, 2005: Proceedings, volume 3381 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-24302-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .S62 2005. URL http:

//www.springerlink.
com/content/978-3-540-
24302-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3381;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105088.

Vassiliadis:2007:ECS

[VBH07] Stamatis Vassiliadis, Mladen Bereković, and Timo D. Hämäläinen, editors. Embedded Computer Systems: Architectures, Modeling, and Simulation: 7th International Workshop, SAMOS 2007, Samos, Greece, July 16– 19, 2007. Proceedings, volume 4599 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73622-0 (print), 3-540-73625-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73625-7.

Vinel:2010:MAC

[VBS^{+}10] Alexey Vinel, Boris Bellalta, Claudio Sacchi, Andrey Lyakhov, Miklós Telek, and Miquel Oliver, editors. Multiple Access Communications: Third International Workshop, MA-COM 2010, Barcelona, Spain, September 13– 14, 2010. Proceedings, volume 6235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15427-1 (print), 3-642-15428-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15428-7.

Ventre:2003:IMN

[VC03] Giorgio Ventre and Roberto Canonico, editors. Interactive Multimedia on Next Generation Networks: First International Workshop on Multimedia Interactive Protocols and Systems, MIPS 2003, Napoli, Italy, November 18– 21, 2003: Proceedings, volume 2899 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20534-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.15 .I6 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2899.
htm;   http://www.springerlink.
com/content/978-3-540-
20534-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2899;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94349.

Vckovski:1999:IGI

[Vck99] Andrej Vckovski, editor. Interoperating geographic information systems: Second International Workshop, INTEROP’99, Zürich, Switzerland, March 10– 12, 1999: Proceedings, volume 1580 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65725-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN G70.212 .I5835 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1580.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65725-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1580.

Visconti:2012:SCN

[VD12] Ivan Visconti and Roberto De Prisco, editors. Security and Cryptography for Networks: 8th International Conference, SCN 2012, Amalfi, Italy, September 5–7, 2012. Proceedings, volume 7485 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32927-6 (print), 3-642-32928-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32928-9.

vanderAalst:2003:ATP

[vdAB03] Wil van der Aalst and Eike Best, editors. Applications and theory of Petri Nets 2003: 24th International Conference, ICATPN 2003, Eindhoven, The Netherlands, June 23–27, 2003: Proceedings, volume 2679 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40334-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A66 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2679.htm;
 http://www.springerlink.
com/content/978-3-540-
40334-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2679.

Also available via the World Wide Web.

vanderAalst:2005:BPM

[vdABCC05] Wil M. P. van der Aalst, Boualem Benatallah, Fabio Casati, and Francisco Curbera, editors. Business process management: 3rd international conference, BPM 2005, Nancy, France, September 5–8, 2005: proceedings, volume 3649 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28238-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.2 .B674 2005. URL

                                        http://springerlink.
                                        metapress.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3649;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        28238-9.

vanderAalst:2000:BPM

[vdADO00] Wil van der Aalst, Jorg Desel, and A. Oberweis, editors. Business Process Management: Models, Techniques, and Empirical Studies, volume 1806 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67454-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 389 pp. LCCN HF5548.2 .B838 2000.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1806.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67454-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1806.

vanderAalst:2003:BPM

[vdAHW03] Wil van der Aalst, Arthur Ter Hofstede, and Mathias Weske, editors. Business process management: International Conference, BPM 2003, Eindhoven, the Netherlands, June 26–27, 2003: Proceedings, volume 2678 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40318-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.2 .B674 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2678.htm;
 http://www.springerlink.
com/content/978-3-540-
40318-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2678.

vandenBrand:2010:SLE

[vdBGG10] Mark van den Brand, Dragan Gašević, and Jeff Gray, editors. Software Language Engineering: Second International Conference, SLE 2009, Denver, CO, USA, October 5–6, 2009, Revised Selected Papers, volume 5969 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12106-3 (print), 3-642-12107-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12107-4.

vandenBerg:2009:WWI

[vdBHOS09] Hans van den Berg, Geert Heijenk, Evgeny Osipov, and Dirk Staehle, editors. Wired/Wireless Internet Communications: 7th International Conference, WWIC 2009, Enschede, The Netherlands, May 27–29, 2009. Proceedings, volume 5546 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02117-4 (print), 3-642-02118-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        02118-3.

vandenBesselaar:2005:DCI

[vdBK05] Peter van den Besselaar and Satoshi Koizumi, editors. Digital Cities III: Information Technologies for Social Capital: Cross-cultural Perspectives. Third International Digital Cities Workshop, Amsterdam, The Netherlands, September 18–19, 2003. Revised Selected Papers, volume 3081 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25331-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.83 .K96 2005.

URL http://www.springerlink.
com/content/978-3-540-
25331-0;   http://www.
springerlink.com/openurl.
asp?genre=volume&id=
doi:10.1007/b107136.

vandenBussche:2001:DTI

[vdBV01] Jan van den Bussche and Victor Vianu, editors. Database theory

— ICDT 2001: 8th International Conference, London, UK, January 4–6, 2001: Proceedings, volume 1973 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41456-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I557 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1973.htm;
 http://www.springerlink.
com/content/978-3-540-
41456-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1973.

Vielhauer:2011:BIM

[VDD^{+}11] Claus Vielhauer, Jana Dittmann, Andrzej Drygajlo, Niels Christian Juul, and Michael C. Fairhurst, editors. Biometrics and ID Management: COST 2101 European Workshop, BioID 2011, Brandenburg (Havel), Germany, March 8– 10, 2011. Proceedings, volume 6583 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-19529-6 (print), 3-642-19530-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        19530-3.

Villa:2012:ANNb

[VDÉ^{+}12a] Alessandro E. Villa, Włodzisław Duch, Péter Érdi, Francesco Masulli, and Günther Palm, editors. Artificial Neural Networks and Machine Learning – ICANN 2012: 22nd International Conference on Artificial Neural Networks, Lausanne, Switzerland, September 11–14, 2012, Proceedings, Part II, volume 7553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33265-X (print), 3-642-33266-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33266-1.

Villa:2012:ANNa

[VDÉ^{+}12b] Alessandro E. P. Villa, Włodzisław Duch, Péter Érdi, Francesco Masulli, and Günther Palm, editors. Artificial Neural Networks and Machine Learning – ICANN 2012: 22nd International Conference on Artificial Neural Networks, Lausanne, Switzerland, September 11–14, 2012, Proceedings, Part I, volume 7552 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33268-4 (print), 3-642-33269-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33269-2.

vandenHerik:2007:CGI

[vdHCD07] H. Jaap van den Herik, Paolo Ciancarini, and H. H. L. M. (Jeroen) Donkers, editors. Computers and Games: 5th International Conference, CG 2006, Turin, Italy, May 29–31, 2006. Revised Papers, volume 4630 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75537-3 (print), 3-540-75538-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75538-8.

vandenHerik:2006:ACG

[vdHHsHD06] H. Jaap van den Herik, Shun-Chin Hsu, Tsan sheng Hsu, and H. H. L. M. (Jeroen) Donkers, editors. Advances in Computer Games: 11th International Conference, ACG 2005, Taipei, Taiwan, September 6–9, 2005. Revised Papers, volume 4250 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48887-1 (print), 3-540-48889-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
48889-7.

vandenHerik:1999:CGF

[vdHI99] Jaap van den Herik and Hiroyuki Iida, editors. Computers and games: First International Conference, CG’98, Tsukuba, Japan, November 11– 12, 1998: Proceedings, volume 1558 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65766-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .C4166 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1558.htm;
http://www.springerlink.
com/content/978-3-540-
65766-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1558.

vandenHerik:2011:CGI

[vdHIP11] H. Jaap van den Herik, Hiroyuki Iida, and Aske Plaat, editors. Computers and Games: 7th International Conference, CG 2010, Kanazawa, Japan, September 24–26, 2010, Revised Selected Papers, volume 6515 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17927-4 (print), 3-642-17928-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17928-0.

vandenHerik:2012:ACG

[vdHP12] H. Jaap van den Herik and Aske Plaat, editors. Advances in Computer Games: 13th International Conference, ACG 2011, Tilburg, The Netherlands, November 20–22, 2011, Revised Selected Papers, volume 7168 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31865-7 (print), 3-642-31866-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31866-5.

vandenHerik:2010:ACG

[vdHS10] H. Jaap van den Herik and Pieter Spronck, editors. Advances in Computer Games: 12th International Conference, ACG 2009, Pamplona Spain, May 11– 13, 2009. Revised Papers, volume 6048 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12992-7 (print), 3-642-12993-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12993-3.

vandenHerik:2008:CGI

[vdHXMW08] H. Jaap van den Herik, Xinhe Xu, Zongmin Ma, and Mark H. M. Winands, editors. Computers and Games: 6th International Conference, CG 2008, Beijing, China, September 29 — October 1, 2008. Proceedings, volume 5131 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87607-3 (print), 3-540-87608-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
87608-3.

vanDam:2011:TQC

[vDKS11] Wim van Dam, Vivien M.

Kendon, and Simone Severini, editors. Theory of Quantum Computation, Communication, and Cryptography: 5th Conference, TQC 2010, Leeds, UK, April 13–15, 2010, Revised Selected Papers, volume 6519 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-18072-8 (print), 3-642-18073-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
18073-6.

vanderLinden:1998:DES

[vdL98] Frank van der Linden, editor. Development and evolution of software architectures for product families / Second International ESPRIT ARES Workshop, Las Palmas de Gran Canaria, Spain, February 26–27, 1998: Proceedings, volume 1429 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64916-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I73 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1429.htm;
 http://www.springerlink.
com/content/978-3-540-
64916-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1429.

vanderLinden:2001:SAP

[vdL01] Frank van der Linden, editor. Software architectures for product families: International Workshop IW-SAPF-3, Las Palmas de Gran Canaria, Spain, March 15–17, 2000: Proceedings, volume 1951 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41480-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I597 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1951.
htm;   http://www.springerlink.
com/content/978-3-540-
41480-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1951.

Contents: Product family practice — Business – Product family concepts

— Product family methods – Evolution — Product family techniques — Surveys.

vanderLinden:2002:SPF

[vdL02] Frank van der Linden, editor. Software product-family engineering: 4th International Workshop, PFE 2001, Bilbao, Spain, October 3–5, 2001: Revised Papers, volume 2290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43659-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 P485 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2290.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43659-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2290.

vanderLinden:2004:SPF

[vdL04] Frank van der Linden, editor. Software Product-Family Engineering: 5th International Workshop, PFE 2003, Siena, Italy, November 4–6, 2003: Revised Papers, volume 3014 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21941-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .L43x v.3014. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3014.
htm;   http://www.springerlink.
com/content/978-3-540-
21941-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3014;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97155.

vanDitmarsch:2011:LRI

[vDLJ11] Hans van Ditmarsch, Jérôme Lang, and Shier Ju, editors. Logic, Rationality, and Interaction: Third International Workshop, LORI 2011, Guangzhou, China, October 10–13, 2011. Proceedings, volume 6953 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24129-8 (print), 3-642-24130-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24130-7.

vanderMeer:2008:MAC

[vdMBD08] Sven van der Meer, Mark Burgess, and Spyros Denazis, editors. Modelling Autonomic Communications Environments: Third IEEE International Workshop, MACE 2008, Samos Island, Greece, September 22–26, 2008. Proceedings, volume 5276 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-87354-6 (print), 3-540-87355-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        87355-6.

vanderMeyden:2011:MCA

[vdMS11] Ron van der Meyden and Jan-Georg Smaus, editors. Model Checking and Artificial Intelligence: 6th International Workshop, MoChArt 2010, Atlanta, GA, USA, July 11, 2010, Revised Selected and Invited Papers, volume 6572 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20673-5 (print), 3-642-20674-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20674-0.

vanderMeyden:2008:DLC

[vdMvdT08] Ron van der Meyden and Leendert van der Torre, editors. Deontic Logic in Computer Science: 9th International Conference, DEON 2008, Luxembourg, Luxembourg, July 15–18, 2008. Proceedings, volume 5076 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70524-4 (print), 3-540-70525-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70525-3.

Ventre:1996:MTA

[VDPD96] G. (Giorgio) Ventre, J. (Jordi) Domingo-Pascual, and A. Danthine, editors. Multimedia telecommunications and applications: Third International COST 237 Workshop, Barcelona, Spain, November 25– 27, 1996: proceedings, volume 1185 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-62096-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I65 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1185.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62096-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1185.

vanderPoorten:2008:ANT

[vdPS08] Alfred J. van der Poorten and Andreas Stein, editors. Algorithmic Number Theory: 8th International Symposium, ANTS-VIII Banff, Canada, May 17–22, 2008 Proceedings, volume 5011 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79455-7 (print), 3-540-79456-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79456-1.

vandePol:2010:MCS

[vdPW10] Jaco van de Pol and Michael Weber, editors. Model Checking Software: 17th International SPIN Workshop, Enschede, The Netherlands, September 27–29, 2010. Proceedings, volume 6349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16163-4 (print), 3-642-16164-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16164-3.

vanderVeer:1984:RCE

[vdV84] G. C. van der Veer, editor. Readings on cognitive ergonomics — mind and computers: proceedings of the 2nd European Conference, Gmunden, Austria, September 10– 14, 1984, volume 178 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 3-540-13394-1 (Berlin), 0-387-13394-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.P75R42 1984. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0178.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        13394-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=178.

Contains papers to be presented at this conference.

vanEijck:1991:LAE

[vE91] J. van Eijck, editor. Logics in AI: European Workshop JELIA ’90, Amsterdam, the Netherlands, September 10–14, 1990, proceedings, volume 478 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-53686-8 (Berlin), 0-387-53686-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .E985 1990.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0478.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        53686-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=478.

vanElst:2004:AMK

[vEDA04] Ludger van Elst, Virginia Dignum, and Andreas Abecker, editors. Agent-Mediated Knowledge Management: International Symposium AMKM 2003, Stanford, CA, USA, March 24–26, 2003: Revised and Invited Papers, volume 2926 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20868-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 A45 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2926.
htm;   http://www.springerlink.
com/content/978-3-540-
20868-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2926;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b94841.

Vatakis:2011:MAT

[VEG^{+}11] Argiro Vatakis, Anna Esposito, Maria Giagkou, Fred Cummins, and Georgios Papadelis, editors. Multidisciplinary Aspects of Time and Time Perception: COST TD0904 International Workshop, Athens, Greece, October 7–8, 2010, Revised Selected Papers, volume 6789 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21477-0 (print), 3-642-21478-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21478-3.

vanEekelen:2011:ITP

[vEGSW11] Marko van Eekelen, Herman Geuvers, Julien Schmaltz, and Freek Wiedijk, editors. Interactive Theorem Proving: Second International Conference, ITP 2011, Berg en Dal, The Netherlands, August 22– 25, 2011. Proceedings, volume 6898 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22862-3 (print), 3-642-22863-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22863-6.

vanEck:2009:AIS

[vEGW09] Pascal van Eck, Jaap Gordijn, and Roel Wieringa, editors. Advanced Information Systems Engineering: 21st International Conference, CAiSE 2009, Amsterdam, The Netherlands, June 8–12, 2009. Proceedings, volume 5565 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02143-3 (print), 3-642-02144-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02144-2.

vanEijk:2005:ACI

[vEHD05] Rogier M. van Eijk, Marc-Philippe Huget, and Frank Dignum, editors. Agent communication: international workshops on agent communication, AC 2004, New York, NY, USA, July 2004, Revised selected and invited papers, volume 3396 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25015-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 I592 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25015-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3396;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b106468.

Veit:2003:MEM

[Vei03] Daniel Veit, editor. Matchmaking in Electronic Markets: an agent-based approach towards matchmaking in electronic negotiations: 2003, volume 2882 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20500-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2882.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20500-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2882;
                                         http://www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b94069.

Veloso:1994:PLA

[Vel94a] Manuela M. Veloso.

Planning and learning by analogical reasoning, volume 886 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58811-6 (Berlin), 0-387-58811-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 181 pp. LCCN Q325.5 .V45 1994. DM46.00.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0886.
htm;   http://www.springerlink.
com/content/978-3-540-
58811-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=886.

Veltkamp:1994:COB

[Vel94b] Remco C. Veltkamp.

Closed object boundaries from scattered points, volume 885 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-58808-6, 3-540-58808-

6. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 144 pp. LCCN T385 .V45 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0885.htm;
 http://www.springerlink.
com/content/978-0-387-
58808-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=885.

“Ph.D. dissertation presents the result of research carried out between 1985 and 1992, first at Leiden University … and later Centre for Mathematics and Computer Science, Amsterdam”–Pref.

Vernon:2000:CVEa

[Ver00a] David Vernon, editor. Computer vision — ECCV 2000: 6th European Conference on Computer Vision, Dublin, Ireland, June 26–July 1, 2000: Proceedings. Part 1, volume 1842 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67685-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .E87 2000; QA267.A1 L43 no.1842-1843. URL

http://link.springer-
ny.com/link/service/

series/0558/tocs/t1842.
htm;   http://www.springerlink.
com/content/978-3-540-
67685-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1842.

Vernon:2000:CVEb

[Ver00b] David Vernon, editor.

Computer vision — ECCV 2000: 6th European Conference on Computer Vision, Dublin, Ireland, June 26–July 1, 2000: Proceedings. Part 2, volume 1843 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67686-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .E87 2000; QA267.A1 L43 no.1842-1843. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1843.
htm;   http://www.springerlink.
com/content/978-3-540-
67686-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1843.

Verdone:2008:WSN

[Ver08] Roberto Verdone, editor. Wireless Sensor Networks: 5th European Conference, EWSN 2008, Bologna, Italy, January 30–February 1, 2008. Proceedings, volume 4913 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77689-3 (print), 3-540-77690-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77690-1.

vanEekelen:2010:FPA

[vES10] Marko van Eekelen and Olha Shkaravska, editors. Foundational and Practical Aspects of Resource Analysis: First International Workshop, FOPARA 2009, Eindhoven, The Netherlands, November 6, 2009, Revised Selected Papers, volume 6324 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15330-5 (print), 3-642-15331-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15331-0.

Vetulani:2011:HLT

[Vet11] Zygmunt Vetulani, editor. Human Language Technology. Challenges for Computer Science and Linguistics: 4th Language and Technology Conference, LTC 2009, Poznan, Poland, November 6–8, 2009, Revised Selected Papers, volume 6562 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20094-X (print), 3-642-20095-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20095-3.

vanEijck:2012:GAS

[vEV12] Jan van Eijck and Rineke Verbrugge, editors. Games, Actions and Social Software: Multidisciplinary Aspects, volume 7010 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29325-5 (print), 3-642-29326-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29326-9.

vanLamsweerde:1991:EES

[vF91] A. van Lamsweerde and A. Fuggetta, editors. ESEC ’91: 3rd European Software Engineering Conference, ESEC ’91, Milan, Italy, October 21–24, 1991: proceedings, volume 550 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54742-8 (Berlin), 0-387-54742-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .E97 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0550.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54742-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=550.

Vazao:2008:INT

[VFC08] Teresa Vazão, Mário M. Freire, and Ilyoung Chong, editors. Information Networking. Towards Ubiquitous Networking and Services: International Conference, ICOIN 2007, Estoril, Portugal, January 23–25, 2007. Revised Selected Papers, volume 5200 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89523-X (print), 3-540-89524-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89524-4.

vanGool:2002:PRD

[vG02] Luc van Gool, editor. Pattern Recognition: 24th DAGM Symposium, Zürich, Switzerland, September 16– 18, 2002. Proceedings, volume 2449 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44209-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1650 .D35 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2449.htm;
 http://www.springerlink.
com/content/978-3-540-
44209-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2449.

Venkatakrishnan:2012:ISS

[VG12] Venkat Venkatakrishnan and Diganta Goswami, editors. Information Systems Security: 8th International Conference, ICISS 2012, Guwahati, India, December 15– 19, 2012. Proceedings, volume 7671 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35129-8 (print), 3-642-35130-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        35130-3.

Vivacqua:2011:CTI

[VGB11] Adriana S. Vivacqua, Carl Gutwin, and Marcos R. S. Borges, editors. Collaboration and Technology: 17th International Conference, CRIWG 2011, Paraty, Brazil, October 2–7, 2011. Proceedings, volume 6969 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23800-9 (print), 3-642-23801-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23801-7.

vonWright:1996:TPH

[vGH96] J. (Joakim) von Wright, J. (Jim) Grundy, and J. (John) Harrison, editors. Theorem proving in higher order logics: 9th International Conference, TPHOLs ’96, Turku, Finland, August 26–30, 1996: proceedings, volume 1125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61587-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96I577 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1125.
htm;   http://www.springerlink.
com/content/978-3-540-
61587-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1125.

Vanneschi:2009:GPE

[VGM^{+}09] Leonardo Vanneschi, Steven Gustafson, Alberto Moraglio, Ivanoe De Falco, and Marc Ebner, editors. Genetic Programming: 12th European Conference, EuroGP 2009 Tübingen, Germany, April 15– 17, 2009 Proceedings, volume 5481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01180-2 (print), 3-642-01181-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01181-8.

Vallecillo:2008:TPM

[VGP08] Antonio Vallecillo, Jeff Gray, and Alfonso Pierantonio, editors. Theory and Practice of Model Transformations: First International Conference, ICMT 2008, Zürich, Switzerland, July 1–2, 2008 Proceedings, volume 5063 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69926-0 (print), 3-540-69927-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69927-9.

vanHulzen:1983:CAE

[vH83] J. A. van Hulzen, editor. Computer algebra: EUROCAL ’83, European Computer Algebra Conference, London, England, March 28– 30, 1983, volume 162 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1983. CO-DEN LNCSD9. ISBN 0-387-12868-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E85 1983. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0162.htm;
 http://www.springerlink.
com/content/978-0-387-
12868-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=162.

Vicente:2004:MMN

[VH04] John Vicente and David Hutchison, editors. Management of multimedia networks and services: 7th IFIP/IEEE international conference, MMNS 2004, San Diego, CA, USA, October 3– 6, 2004: proceedings, volume 3271 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23239-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.15 .I58 2004. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        23239-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3271;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b101051.

vanHemert:2008:ECC

[vHC08] Jano van Hemert and Carlos Cotta, editors. Evolutionary Computation in Combinatorial Optimization: 8th European Conference, Evo-COP 2008, Naples, Italy, March 26–28, 2008. Proceedings, volume 4972 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78603-1 (print), 3-540-78604-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        78604-7.

vanHoeve:2009:IAT

[vHH09] Willem-Jan van Hoeve and John N. Hooker, editors. Integration of AI and OR Techniques in Constraint Programming for Combinatorial Optimization Problems: 6th International Conference, CPAIOR 2009 Pittsburgh, PA, USA, May 27–31, 2009 Proceedings, volume 5547 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01928-5 (print), 3-642-01929-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01929-6.

Vidal:2007:DVI

[VHM07] René Vidal, Anders Heyden, and Yi Ma, editors. Dynamical Vision: ICCV 2005 and ECCV 2006 Workshops, WDV 2005 and WDV 2006, Beijing, China, October 21, 2005, Graz, Austria, May 13, 2006. Revised Papers, volume 4358 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70931-2 (print), 3-540-70932-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        70932-9.

vanHee:2008:ATP

[vHV08] Kees M. van Hee and Rüdiger Valk, editors. Applications and Theory of Petri Nets: 29th International Conference, PETRI NETS 2008, Xi’an, China, June 23– 27, 2008. Proceedings, volume 5062 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68745-9 (print), 3-540-68746-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68746-7.

Vidal:2012:LBP

[Vid12] Germán Vidal, editor. Logic-Based Program Synthesis and Transformation: 21st International Symposium, LOP-STR 2011, Odense, Denmark, July 18–20, 2011. Revised Selected Papers, volume 7225 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32210-7 (print), 3-642-32211-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32211-2.

Vigna:1998:MAS

[Vig98] Giovanni Vigna, editor.

Mobile Agents and Security, volume 1419 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64792-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 256 pp. LCCN QA76.76.I58 M63 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1419.htm;
http://www.springerlink.
com/content/978-3-540-
64792-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1419.

Visser:2004:III

[Vis04] Ubbo Visser. Intelligent information integration for the Semantic Web, volume 3159 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22993-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
22993-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3159;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100348.

Vitanyi:1995:CLT

[Vit95] P. M. B. Vitanyi, editor. Computational learning theory: second European conference, EuroCOLT ’95, Barcelona, Spain, March 13–15, 1995: proceedings, volume 904 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-59119-2 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.7 .E973 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0904.htm;
 http://www.springerlink.
com/content/978-3-540-
59119-1;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=904.

Vitek:2008:EOO

[Vit08] Jan Vitek, editor. ECOOP 2008 — Object-Oriented Programming: 22nd European Conference Paphos, Cyprus, July 7– 11, 2008 Proceedings, volume 5142 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70591-0 (print), 3-540-70592-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70592-5.

Vitek:2010:OMC

[Vit10] Jan Vitek, editor. Objects, Models, Components, Patterns: 48th International Conference, TOOLS 2010, Málaga, Spain, June 28–July 2, 2010. Proceedings, volume 6141 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13952-3 (print), 3-642-13953-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13953-6.

Vitek:1999:SIP

[VJ99] Jan Vitek and Christian D. Jensen. Secure Internet programming: security issues for mobile and distributed objects, volume 1603 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66130-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 501 pp. LCCN QA76.625 .S43 1999.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1603.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        66130-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1603.

Valtchev:2011:FCA

[VJ11] Petko Valtchev and Robert Jäschke, editors. Formal Concept Analysis: 9th International Conference, ICFCA 2011, Nicosia, Cyprus, May 2–6, 2011. Proceedings, volume 6628 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20513-5 (print), 3-642-20514-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20514-9.

Veidenbaum:2003:HPC

[VJAA03] Alex Veidenbaum, Kazuki Joe, Hideharu Amano, and Hideo Aiso, editors. High Performance Computing: 5th International Symposium, ISHPC 2003, Tokyo-Odaiba, Japan, October 20–22, 2003: Proceedings, volume 2858 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20359-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I5855 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2858.htm;
 http://www.springerlink.
com/content/978-3-540-
20359-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2858;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b14207.

Valero:2000:HPCa

[VJKT00] Mateo Valero, Kazuki Joe, Masaru Kitsuregawa, and Hidehiko Tanaka, editors. High Performance Computing: Third International Symposium, ISHPC 2000, Tokyo, Japan, October 16–18, 2000. Proceedings, volume 1940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41128-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1940. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1940.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41128-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1940.

Vrancx:2012:ALA

[VKG12] Peter Vrancx, Matthew Knudson, and Marek Grześ, editors. Adaptive and Learning Agents: International Workshop, ALA 2011, Held at AA-MAS 2011, Taipei, Taiwan, May 2, 2011, Revised Selected Papers, volume 7113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28498-1 (print), 3-642-28499-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28499-1.

Vigna:2003:RAI

[VKJ03] Giovanni Vigna, Christopher Kruegel, and Erland Jonsson, editors. Recent Advances in Intrusion Detection: 6th International Symposium, RAID 2003, Pittsburgh, PA, USA, September 8–10, 2003: Proceedings, volume 2820 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40878-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 R34 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2820.
htm;   http://www.springerlink.
com/content/978-3-540-
40878-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2820;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b13476.

Vilhjalmsson:2011:IVA

[VKMT11] Hannes Högni Vilhjálmsson, Stefan Kopp, Stacy Marsella, and Kristinn R. Thórisson, editors. Intelligent Virtual Agents: 10th International Conference, IVA 2011, Reykjavik, Iceland, September 15–17, 2011. Proceedings, volume 6895 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23973-0 (print), 3-642-23974-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        23974-8.

vanKreveld:2012:GDI

[vKS12] Marc van Kreveld and Bettina Speckmann, editors. Graph Drawing: 19th International Symposium, GD 2011, Eindhoven, The Netherlands, September 21–23, 2011, Revised Selected Papers, volume 7034 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25877-8 (print), 3-642-25878-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25878-7.

Valero:2000:HPCb

[VKV00] Mateo Valero, Viktor K. Prasanna Kumar, and Sriram Vajapeyam, editors. High performance computing — HiPC 2000: 7th International Conference, Bangalore, India, December 17–20, 2000: Proceedings, volume 1970 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41429-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1970. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1970.htm;
http://www.springerlink.
com/content/978-3-540-
41429-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1970.

vanLeeuwen:1988:DAI

[vL88] J. van Leeuwen, editor. Distributed algorithms: 2nd International Workshop, Amsterdam, The Netherlands, July 8–10, 1987: proceedings, volume 312 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 0-387-19366-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 D4851 1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0312.htm;
 http://www.springerlink.
com/content/978-0-387-
19366-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=312.

“The 2nd International Workshop on Distributed Algorithms was held at the Centre for Mathematics and Computer Science (CWI) in Amsterdam, July 8–10, 1987”–Pref.

vanLeeuwen:1989:GTC

[vL89] J. van Leeuwen, editor. Graph-theoretic concepts in computer science: International Workshop WG ’88, Amsterdam, The Netherlands, June 15–17, 1988: proceedings, volume 344 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-50728-0 (USA) (W. Ger.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA166 .G71721 1989. DM60.50. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0344.htm;
 http://www.springerlink.
com/content/978-0-387-
50728-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=344.

The 14th International Workshop on Graph-Theoretic Concepts in Computer Science (WG’88) was held at the Centre for Mathematics and Computer Science (CWI) in Amsterdam, June 15– 17, 1988.

vanLeeuwen:1994:AES

[vL94a] J. van Leeuwen, editor.

Algorithms — ESA ’94: second annual European symposium, Utrecht, The Netherlands, September 26–28, 1994: proceedings, volume 855 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58434-X (Berlin), 0-387-58434-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 E83 1994. DM96.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0855.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        58434-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=855.

vanLeeuwen:1994:GTC

[vL94b] Jan van Leeuwen, editor. Graph-theoretic concepts in computer science: 19th International Workshop, WG ’93, Utrecht, The Netherlands, June 16–18, 1993: proceedings, volume 790 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-86636-292-4 (New York), 3-540-57899-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I647 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0790.htm;
                                         http://www.springerlink.
                                        com/content/978-0-86636-
                                        292-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=790.

vanLeeuwen:1995:CST

[vL95] J. van Leeuwen. Computer science today: recent trends and developments, volume 1000 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60105-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 641 pp. LCCN QA76. .C57326 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1000.
htm;   http://www.springerlink.
com/content/978-3-540-
60105-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1000.

vanLeeuwen:1980:ALP

[vLdB80] J. van Leeuwen and J. W. de Bakker, editors. Automata, languages, and programming: seventh colloquium, Noordwijkerhout, the Netherlands, July 14– 18, 1980, volume 85 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CO-DEN LNCSD9. ISBN 0-387-10003-2 (paperback: New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.85.

URL http://link.springer-
ny.com/link/service/

                                        series/0558/tocs/t0085.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        10003-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=85.

Sponsored by the European Association for Theoretical Computer Science.

vanLeeuwen:2007:STP

[vLIvdH^{+}07] Jan van Leeuwen, Giuseppe F. Italiano, Wiebe van der Hoek, Christoph Meinel, Harald Sack, and František Plášil, editors. SOFSEM 2007: Theory and Practice of Computer Science: 33rd Conference on Current Trends in Theory and Practice of Computer Science, Harrachov, Czech Republic, January 20–26, 2007. Proceedings, volume 4362 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69506-0 (print), 3-540-69507-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69507-3.

vanLeeuwen:2010:STP

[vLMP^{+}10] Jan van Leeuwen, Anca Muscholl, David Peleg, Jaroslav Pokorný, and Bernhard Rumpe, editors. SOFSEM 2010: Theory and Practice of Computer Science: 36th Conference on Current Trends in Theory and Practice of Computer Science, Špindlerův Mlýn, Czech Republic, January 23–29, 2010. Proceedings, volume 5901 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11265-X (print), 3-642-11266-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11266-9.

vanLeeuwen:1991:DAI

[vLS91] J. van Leeuwen and N. Santoro, editors. Distributed algorithms: 4th International Workshop, Bari, Italy, September 24–26, 1990: proceedings, volume 486 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-54099-7 (Berlin), 0-387-54099-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D565 1991. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0486.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        54099-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=486.

“The 4th International Workshop on Distributed Algorithms (WDAG 90) was organized by the Istituto di scienze dell’informazione of the University of Bari (Italy) and held in Serra Alimini (near Otranto), some 200 kilometers southeast of Bari along the beautiful Adriatic coast”–Pref.

vanLeuken:2011:ICS

[vLS11] René van Leuken and Gilles Sicard, editors. Integrated Circuit and System Design. Power and Timing Modeling, Optimization, and Simulation: 20th International Workshop, PAT-MOS 2010, Grenoble, France, September 7–10, 2010, Revised Selected Papers, volume 6448 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-17751-4 (print), 3-642-17752-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17752-1.

vanLeeuwen:2000:TCS

[vLWH^{+}00] J. van Leeuwen, O. Watanabe, M. Hagiya, P. D. Mosses, and T. Ito, editors. Theoretical computer science: exploring new frontiers of theoretical informatics: International Conference IFIP TCS 2000, Sendai, Japan, August 17–19, 2000: Proceedings, volume 1872 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67823-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I54446 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1872.htm;
http://www.springerlink.
com/content/978-3-540-
67823-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1872.

Vossen:2009:WIS

[VLY09] Gottfried Vossen, Darrell D. E. Long, and Jeffrey Xu Yu, editors. Web Information Systems Engineering — WISE 2009:

10th International Conference, Poznań, Poland, October 5–7, 2009. Proceedings, volume 5802 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04408-5 (print), 3-642-04409-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04409-0.

vonLuck:1994:MPC

[vM94] Kai von Luck and H. Marburger, editors. Management and processing of complex data structures: Third Workshop on Information Systems and Artificial Intelligence, Hamburg, Germany, February 28– March 2, 1994: proceedings, volume 777 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 0-387-57802-1 (New York), 3-540-57802-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D35 W667 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0777.htm;

http://www.springerlink.
com/content/978-0-387-
57802-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=777.

Varadharajan:1999:ICS

[VM99] Vijay Varadharajan and Yi Mu, editors. Information and communication security: Second International Conference, ICICS’99, Sydney, Australia, November 9–11, 1999: Proceedings, volume 1726 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66682-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I554 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1726.htm;
http://www.springerlink.
com/content/978-3-540-
66682-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1726.

vanMelkebeek:2000:RCC

[vM00] Dieter van Melkebeek.

Randomness and Completeness in Computational Complexity, volume 1950 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41492-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 196 pp. LCCN QA76.9.M35 M54 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1950.htm;
 http://www.springerlink.
com/content/978-3-540-
41492-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1950.

Revision of thesis (Ph. D.)–University of Chicago, 1999. Contents: Preliminaries — Derandomizing Arthur–Merlin games

— Sparseness of complete languages — Autoreducibility of complete languages — Size of randomized polynomial time — Frequency of complete languages

— Frequency of antoreducible languages.

Varadharajan:2001:ISP

[VM01] Vijay Varadharajan and Yi Mu, editors. Information security and privacy: 6th Australasian Conference, ACISP 2001, Sydney, Australia, July 11–13, 2001: Proceedings, volume 2119 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42300-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Electronic book; QA76.9.A25 A278 2001; QA267.A1 L43 no.2119. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2119.htm;
 http://www.springerlink.
com/content/978-3-540-
42300-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2119.

Vinel:2012:CTV

[VMB^{+}12] Alexey Vinel, Rashid Mehmood, Marion Berbineau, Cristina Rico Garcia, Chung-Ming Huang, and Naveen Chilamkurti, editors. Communication Technologies for Vehicles: 4th International Workshop, Nets4Cars/Nets4Trains 2012, Vilnius, Lithuania, April 25–27, 2012. Proceedings, volume 7266 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29666-1 (print), 3-642-29667-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29667-3.

Vicedo:2004:ANL

[VMBMN04] José Luis Vicedo, Patricio Martànez-Barco, Rafael Muñoz, and Maximiliano Saiz Noeda, editors. Advances in Natural Language Processing: 4th International Conference, ESTAL 2004, Alicante, Spain, October 20–22, 2004, Proceedings, volume 3230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23498-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 F67. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3230.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        23498-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3230;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b101638.

Varbanescu:2012:CAI

[VMvN12] Ana Lucia Varbanescu, Anca Molnos, and Rob van Nieuwpoort, editors. Computer Architecture: ISCA 2010 International Workshops A4MMC, AMAS-BT, EAMA, WEED, WIOSCA, Saint-Malo, France, June 19–23, 2010, Revised Selected Papers, volume 6161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24321-5 (print), 3-642-24322-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24322-6.

vanOostrom:2004:RTA

[vO04] Vincent van Oostrom, editor. Rewriting Techniques and Applications: 15th International Conference, RTA 2004, Aachen, Germany, June 3–5, 2004, Proceedings, volume 3091 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22153-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN x + 311. URL http:

//link.springer-ny.

                                        com/link/service/series/
                                        0558/tocs/t3091.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        22153-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3091;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98160.

Vogt:1988:CIC

[Vog88] F. H. Vogt, editor. Concurrency 88, International Conference on Concurrency, Hamburg, FRG, October 18– 19, 1988: proceedings, volume 335 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CODEN LNCSD9. ISBN 0-387-50403-6 (USA) (Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .C61851 1988. DM55.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0335.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        50403-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=335.

Vogler:1992:MCP

[Vog92] W. (Walter) Vogler.

Modular construction and partial order semantics of Petri nets, volume 625 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55767-9 (Berlin), 0-387-55767-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 252 pp. LCCN QA76.58 .V63 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0625.htm;
 http://www.springerlink.
com/content/978-3-540-
55767-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=625.

Voges:2001:CSR

[Vog01] U. Voges, editor. Computer safety, reliability and security: 20th International Conference, SAFECOMP 2001, Budapest, Hungary, September 26–28, 2001: Proceedings, volume 2187 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42607-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2187. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2187.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        42607-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2187.

Volkert:1993:PCS

[Vol93] Jens Volkert, editor. Parallel computation: Second International ACPC Conference, Gmunden, Austria, October 4– 6, 1993: proceedings, volume 734 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-57314-3 (U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5 1993. DM58.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0734.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        57314-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=734.

vonBochmann:1979:ADC

[von79] Gregor von Bochmann. Architecture of distributed computer systems, volume 77 of Lecture Notes in Computer Science.

Springer-Verlag Inc., New York, NY, USA, 1979. CODEN LNCSD9. ISBN 0-387-09723-6. ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 238 pp. LCCN QA76.9.D5 .B63. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0077.
htm;   http://www.springerlink.
com/content/978-0-387-
09723-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=77.

vonMartial:1992:CPA

[von92] F. (Frank) von Martial. Coordinating plans of autonomous agents, volume 610 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 0-387-55615-X (New York), 3-540-55615-X (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 246 pp. LCCN Q337 .M36 1992. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0610.
htm;   http://www.springerlink.
com/content/978-0-387-
55615-4;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=610.

vonderMalsburg:1996:ANN

[von96] Christoph von der Malsburg, editor. Artificial neural networks: ICANN 96: 6th international conference, Bochum, Germany, July 16–19, 1996: proceedings, volume 1112 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61510-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .A74335 1996. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1112.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61510-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1112.

Voronkov:1992:LPA

[Vor92a] A. Voronkov, editor. Logic programming and automated reasoning: international conference, LPAR ’92, St. Petersburg, Russia, July 15– 20, 1992: proceedings, volume 624 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55727-X (Berlin), 0-387-55727-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I55 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0624.htm;
 http://www.springerlink.
com/content/978-3-540-
55727-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=624.

Voronkov:1992:LPF

[Vor92b] A. Voronkov, editor.

Logic programming: First Russian Conference on Logic Programming, Irkutsk, Russia, September 14– 18, 1990: Second Russian Conference on Logic Programming, St. Petersburg, Russia, September 11–16, 1991: proceedings, volume 592 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55460-2 (Berlin), 0-387-55460-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .R87 1990. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0592.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55460-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=592.

Voronkov:1993:LPA

[Vor93] A. Voronkov, editor. Logic programming and automated reasoning: 4th international conference, LPAR ’93, St. Petersburg, Russia, July 13– 20, 1993: proceedings, volume 698 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56944-8 (Berlin), 0-387-56944-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I55 1993. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0698.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        56944-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=698.

Voronkov:2002:ADC

[Vor02] A. Voronkov, editor.

Automated Deduction— CADE-18: 18th International Conference on Automated Deduction, Copenhagen, Denmark, July 27–30, 2002. Proceedings, volume 2392 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43931-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 I57 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2392.htm;
 http://www.springerlink.
com/content/978-3-540-
43931-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2392.

Voronkov:2008:RTA

[Vor08] Andrei Voronkov, editor. Rewriting Techniques and Applications: 19th International Conference, RTA 2008 Hagenberg, Austria, July 15–17, 2008 Proceedings, volume 5117 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70588-0 (print), 3-540-70590-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70590-1.

Vosselman:1992:RM

[Vos92] G. (George) Vosselman. Relational matching, volume 628 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 0-387-55798-9 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 190 pp. LCCN TA1632 .V6715 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0628.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        55798-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=628.

Thesis (doctoral) — Rheinische Friedrich-Wilhelms-Universität, Bonn.

Voss:2003:OSM

[Vos03] Michael J. Voss, editor. OpenMP shared memory parallel programming: International Workshop on OpenMP Applications and Tools, WOMPAT 2003, Toronto, Canada, June 26–27, 2003: Proceedings, volume 2716 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40435-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .I589 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2716.htm;
 http://www.springerlink.
com/content/978-3-540-
40435-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2716.

vonBochmann:1993:CAV

[vP93] Gregor von Bochmann and D. K. Probst, editors. Computer aided verification: fourth international workshop, CAV ’92, Montréal, Canada, June 29–July 1, 1992: proceedings, volume 663 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56496-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E93 C38 1992.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0663.
htm;   http://www.springerlink.
com/content/978-0-387-
56496-8;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=663.

VandeVelde:1996:ABE

[VP96] Walter Van de Velde and John W. Perram, editors. Agents breaking away: 7th European Workshop on Modelling Autonomous Agents in a Multi-Agent World, MAAMAW ’96, Eindhoven, Netherlands, January 22–25, 1996: proceedings, volume 1038 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-60852-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 E78 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1038.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        60852-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1038.

Vouros:2004:MAA

[VP04] George A. Vouros and Themistoklis Panayiotopoulos, editors. Methods and Applications of Artificial Intelligence: Third Hellenic conference on AI, SETN 2004: Samos, Greece, May 5–8 2004: Proceedings, volume 3025 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21937-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3025.htm;
 http://www.springerlink.
com/content/978-3-540-
21937-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3025;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97168.

VanEmdeBoas:2003:STP

[VPBŠ03] Peter Van Emde Boas, Jaroslav Pokorný, Mária Bieliková, and Július Štuller, editors. SOF-SEM 2004: Theory and Practice of Computer Science: Theory and Practice of Computer Science: 30th Conference on Current Trends in Theory and Practice of Computer Science, Měřín, Czech Republic, January 24–30, 2004: Proceedings, volume 2932 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20779-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .S62 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2932.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20779-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2932;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95046.

Veloso:2000:RRS

[VPK00] Manuela M. Veloso, Enrico Pagello, and Hiroaki Kitano, editors. RoboCup-99: Robo Soccer World Cup III, volume 1856 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-41043-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ210.3 .R6153 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1856.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-

41043-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1856.

Varadharajan:1997:ISP

[VPM97] Vijay Varadharajan, Josef Pieprzyk, and Yi Mu, editors. Information security and privacy: second Australasian conference, ACISP’97, Sydney, NSW, Australia, July 7–9, 1997: proceedings, volume 1270 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63232-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN A76.9.A25A279 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1270.htm;
http://www.springerlink.
com/content/978-3-540-
63232-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1270.

Velasquez:2009:KBIa

[VRHJ09a] Juan D. Velásquez, Sebastián A. Ríos, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 13th International Conference, KES 2009, Santiago, Chile, September 28–30, 2009, Proceedings, Part I, volume 5711 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04594-4 (print), 3-642-04595-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04595-0.

Velasquez:2009:KBIb

[VRHJ09b] Juan D. Velásquez, Sebastián A. Ríos, Robert J. Howlett, and Lakhmi C. Jain, editors. Knowledge-Based and Intelligent Information and Engineering Systems: 13th International Conference, KES 2009, Santiago, Chile, September 28–30, 2009, Proceedings, Part II, volume 5712 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04591-X (print), 3-642-04592-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04592-9.

Visser:2008:RRS

[VROD08] Ubbo Visser, Fernando Ribeiro, Takeshi Ohashi, and Frank Dellaert, editors. RoboCup 2007: Robot Soccer World Cup XI, volume 5001 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68846-3 (print), 3-540-68847-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68847-1.

Vlahavas:2002:MAA

[VS02] Ioannis Vlahavas and Constantine D. Spyropoulos, editors. Methods and applications of artificial intelligence: second Hellenic Conference on AI, SETN 2002, Thessaloniki, Greece, April 11–12, 2002: Proceedings, volume 2308 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43472-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .H45 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2308.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        43472-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2308.

Voelker:2005:PPS

[VS05] Geoffrey M. Voelker and Scott Shenker, editors. Peer-to-Peer Systems III: Third International Workshop, IPTPS 2004, La Jolla, CA, USA, February 26–27, 2004, Revised Selected Papers, volume 3279 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24252-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24252-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=volume&id=
                                        doi:10.1007/b104020.

Valadas:2009:TMT

[VS09] Rui Valadas and Paulo Salvador, editors. Traffic Management and Traffic Engineering for the Future Internet: First Euro-NF Workshop, FITraMEn 2008, Porto, Portugal, December 11–12, Revised Selected Papers, volume 5464 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04575-8 (print), 3-642-04576-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04576-9.

vanSteen:2006:MAI

[vSH06] Maarten van Steen and Michi Henning, editors. Middleware 2006: ACM/ IFIP/USENIX 7th International Middleware Conference, Melbourne, Australia, November 27– December 1, 2006. Proceedings, volume 4290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49023-X (print), 3-540-68256-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68256-1.

Vitria:2011:PRI

[VSH11] Jordi Vitrià, João Miguel Sanches, and Mario Hernández, editors. Pattern Recognition and Image Analysis: 5th Iberian Conference, IbPRIA 2011, Las Palmas de Gran Canaria, Spain, June 8– 10, 2011. Proceedings, volume 6669 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21256-5 (print), 3-642-21257-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21257-4.

Villatoro:2012:MAB

[VSMS12] Daniel Villatoro, Jordi Sabater-Mir, and Jaime Simão Sichman, editors. Multi-Agent-Based Simulation XII: International Workshop, MABS 2011, Taipei, Taiwan, May 2–6, 2011, Revised Selected Papers, volume 7124 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28399-3 (print), 3-642-28400-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
28400-7.

vanSinderen:2001:PMS

[vSN01] Marten J. van Sinderen and Lambert J. M. Nieuwenhuis, editors. Protocols for multimedia systems: 6th International Conference, PROMS 2001, Enschede, the Netherlands, October 17–19, 2001: Proceedings, volume 2213 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42708-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .P76 2001; QA267.A1 L43 no.2213. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2213.htm;
http://www.springerlink.
com/content/978-3-540-
42708-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2213.

Vogt:2006:SGB

[VSTN06] Paul Vogt, Yuuya Sugita, Elio Tuci, and Chrystopher Nehaniv, editors. Symbol Grounding and Beyond: Third International Workshop on the Emergence and Evolution of Linguistic Communication, EELC 2006, Rome, Italy, September 30 – October 1, 2006. Proceedings, volume 4211 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45769-0 (print), 3-540-45771-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45771-8.

Vitek:1997:MOS

[VT97] Jan Vitek and Christian Tschudin, editors. Mobile object systems: towards the programmable Internet: second International Workshop, MOS ’96, Linz, Austria, July 8–9, 1996: selected presentations and invited papers, volume 1222 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62852-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.625 .M68 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1222.htm;

http://www.springerlink.
com/content/978-3-540-
62852-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1222.

Vallecillo:2012:MFA

[VTK^{+}12] Antonio Vallecillo, Juha-Pekka Tolvanen, Ekkart Kindler, Harald Störrle, and Dimitris Kolovos, editors. Modelling Foundations and Applications: 8th European Conference, ECMFA 2012, Kgs. Lyngby, Denmark, July 2–5, 2012. Proceedings, volume 7349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31490-2 (print), 3-642-31491-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31491-9.

Vene:2005:AFP

[VU05] Varmo Vene and Tarmo Uustalu, editors. Advanced functional programming: 5th international school, AFP 2004, Tartu, Estonia, August 14–21, 2004, revised lectures, volume 3622 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28540-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28540-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3622.

Vetulani:2009:HLT

[VU09] Zygmunt Vetulani and Hans Uszkoreit, editors. Human Language Technology. Challenges of the Information Society: Third Language and Technology Conference, LTC 2007, Poznan, Poland, October 5– 7, 2007, Revised Selected Papers, volume 5603 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04234-1 (print), 3-642-04235-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04235-5.

Vardi:2003:LPA

[VV03] Moshe Y. Vardi and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 10th International Conference, LPAR 2003, Almaty, Kazakhstan, September 22–26, 2003: Proceedings, volume 2850 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20101-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .L735 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2850.htm;
 http://www.springerlink.
com/content/978-3-540-
20101-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2850;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13986.

Virbitskaite:2007:PSI

[VV07] Irina Virbitskaite and Andrei Voronkov, editors. Perspectives of Systems Informatics: 6th International Andrei Ershov Memorial Conference, PSI 2006, Novosibirsk, Russia, June 27– 30, 2006. Revised Papers, volume 4378 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-70880-4 (print), 3-540-70881-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        70881-0.

Vaandrager:1999:HSC

[VvS99] Frits Vaandrager and J. H. van Schuppen, editors. Hybrid systems: computation and control: Second International Workshop, HSCC’99, Berg en Dal, The Netherlands, March 29–31, 1999: Proceedings, volume 1569 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65734-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.38.H73 1999.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1569.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65734-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1569.

vanSomeren:1997:MLE

[vW97] Maarten W. van Someren and Gerhard Widmer, editors. Machine learning: ECML-97: 9th European Conference on Machine Learning, Prague, Czech Republic, April 23–25, 1997: proceedings, volume 1224 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62858-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1224.htm;
 http://www.springerlink.
com/content/978-3-540-
62858-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1224.

Vardanega:2005:RST

[VW05] Tullio Vardanega and Andy Wellings, editors. Reliable Software Technology — Ada-Europe 2005: 10th Ada-Europe International Conference on Reliable Software Technologies, York, UK, June 20–24, 2005. Proceedings, volume 3555 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-26286-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 A33 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        26286-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3555.

VanHentenryck:2007:IAT

[VW07] Pascal Van Hentenryck and Laurence Wolsey, editors. Integration of AI and OR Techniques in Constraint Programming for Combinatorial Optimization Problems: 4th International Conference, CPAIOR 2007, Brussels, Belgium, May 23–26, 2007. Proceedings, volume 4510 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72396-X (print), 3-540-72397-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        72397-4.

Vingron:2008:RCM

[VW08] Martin Vingron and Limsoon Wong, editors. Research in Computational Molecular Biology: 12th Annual International Conference, RECOMB 2008, Singapore, March 30 — April 2, 2008. Proceedings, volume 4955 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78838-7 (print), 3-540-78839-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78839-3.

Vassiliadis:2006:ECS

[VWH06] Stamatis Vassiliadis, Stephan Wong, and Timo D. Hämäläinen, editors. Embedded Computer Systems: Architectures, Modeling, and Simulation: 6th International Workshop, SAMOS 2006, Samos, Greece, July 17– 20, 2006. Proceedings, volume 4017 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36410-2 (print), 3-540-36411-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        36411-5.

vanEmdeBoas:1999:ALP

[vWN99] P. van Emde Boas, J. Wiedermann, and M. Nielsen, editors. Automata, languages and programming: 26th international colloquium, ICALP’99, Prague, Czech Republic, July 11–15, 1999: Proceedings, volume 1644 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66224-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I23 1999. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1644.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66224-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1644.

Vulkov:1997:NAA

[VWY97] Lubin Vulkov, Jerzy Waśniewski, and Plamen Yalamov, editors. Numerical analysis and its applications: first International Workshop, WNAA ’96, Rousse, Bulgaria, June 24–26, 1996: proceedings, volume 1196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62598-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA297 .W584 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1196.htm;
 http://www.springerlink.
com/content/978-3-540-
62598-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1196.

Vulkov:2001:NAA

[VWY01] Lubin Vulkov, Jerzy Waśniewski, and Plamen Yalamov, editors. Numerical analysis and its applications: Second International Conference, NAA 2000, Rousse, Bulgaria, June 11–15, 2000: Revised Papers, volume 1988 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41814-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA297 .N8415 2001; QA267.A1 L43 no.1988. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1988.htm;

                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41814-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1988.

Valverde:1993:IAM

[VY93] Llorenc Valverde and Ronald R. Yager, editors. IPMU ’92, advanced methods in artificial intelligence: 4th International Conference on Information Processing and Management of Uncertainty in Knowledge-Based Systems, Palma de Mallorca, Spain, July 6–10, 1992: proceedings, volume 682 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-56735-6 (New York), 3-540-56735-6 (Berlin: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .I557 1992. US$54.00, DM72.00. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0682.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        56735-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=682.

Vaudenay:2001:SAC

[VY01] Serge Vaudenay and Amr M. Youssef, editors. Selected Areas in Cryptography: 8th Annual International Workshop, SAC 2001 Toronto, Ontario, Canada, August 16–17, 2001. Revised Papers, volume 2259 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-43066-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 S22 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2259.
htm;   http://www.springerlink.
com/content/978-3-540-
43066-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2259.

Vytopil:1991:FTR

[Vyt91] J. Vytopil, editor. Formal techniques in realtime and fault-tolerant systems: Second International Symposium, Nijmegen, the Netherlands, January 8–10, 1992, proceedings, volume 571 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-55092-5 (Berlin), 0-387-55092-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.54 .F66 1991. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0571.htm;
 http://www.springerlink.
com/content/978-3-540-
55092-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=571.

Vitter:1999:AEI

[VZ99] Jeffrey Scott Vitter and Christos D. Zaroliagis, editors. Algorithm engineering: 3rd International Workshop, WAE’99 London, UK, July 19–21, 1999: Proceedings, volume 1668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66427-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1668; QA76.9.A43 I59 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1668.
htm;   http://www.springerlink.
com/content/978-3-540-
66427-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1668.

Valdes:2006:RAI

[VZ06] Alfonso Valdes and Diego Zamboni, editors. Recent advances in intrusion detection: 8th international symposium, RAID 2005, Seattle, WA, USA, September 7–9, 2005, revised papers, volume 3858 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31779-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
31779-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3858.

vonzurGathen:2008:AFF

[vzGIK08] Joachim von zur Gathen, José Luis Imaña, and Çetin Kaya Koç, editors. Arithmetic of Finite Fields: 2nd International Workshop, WAIFI 2008 Siena, Italy, July 6–9, 2008 Proceedings, volume 5130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69498-6 (print), 3-540-69499-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69499-1.

Vigneron:2010:LVA

[VZM^{+}10] Vincent Vigneron, Vicente Zarzoso, Eric Moreau, Rémi Gribonval, and Emmanuel Vincent, editors. Latent Variable Analysis and Signal Separation: 9th International Conference, LVA/ICA 2010, St. Malo, France, September 27–30, 2010. Proceedings, volume 6365 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15994-X (print), 3-642-15995-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15995-4.

vanZee:1989:PC

[vZvdV89] G. A. van Zee and J. G. G. van de Vorst, editors. Parallel computing, 1988, volume 384 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51604-2 (West Germany). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S539 1988. DM28.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0384.htm;
http://www.springerlink.
com/content/978-0-387-
51604-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=384.

Wang:2012:WIS

[W^{+}12] Fu Lee Wang et al., editors. Web information systems and mining: international conference, WISM 2012, Chengdu, China, October 26–28, 2012, proceedings, volume 7529 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-33468-7, 3-642-33469-5 (e-book). LCCN QA76.9.D343 W57 2012.

URL http://www.springerlink.
com/content/978-3-642-
33468-9;   http://www.
springerlink.com/content/
978-3-642-33469-6.

Wada:1986:LPP

[Wad86] Eiiti Wada, editor. Logic programming ’85: proceedings of the 4th conference, Tokyo, Japan, July 1–3, 1985, volume 221 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CODEN LNCSD9. ISBN 0-387-16479-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L58751 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0221.htm;
 http://www.springerlink.
com/content/978-0-387-
16479-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=221.

Papers presented at the Fourth Logic Programming Conference.

Wada:1987:LPP

[Wad87] Eiiti Wada, editor. Logic programming ’86: proceedings of the 5th Conference, Tokyo, Japan, June 23–26, 1986, volume 264 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CODEN LNCSD9. ISBN 0-387-18024-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L58761 1987. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t0264.htm;
 http://www.springerlink.
com/content/978-0-387-
18024-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=264.

“Papers presented at the Fifth Logic Programming Conference … sponsored by ICOT (the Institute for New Generation Computer Technology)”–Frwd.

Wagner:1994:VLK

[Wag94] G. (Gerd) Wagner. Vivid logic: knowledge-based reasoning with two kinds of negation, volume 764 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-57604-5 (Berlin), 0-387-57604-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 148 pp. LCCN QA76.76.E95 W33 1994.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0764.
htm;   http://www.springerlink.
com/content/978-3-540-
57604-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=764.

Wagner:2008:ACC

[Wag08] David Wagner, editor. Advances in Cryptology — CRYPTO 2008: 28th Annual International Cryptology Conference, Santa Barbara, CA, USA, August 17– 21, 2008. Proceedings, volume 5157 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85173-9 (print), 3-540-85174-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85174-5.

Wallach:1982:ASP

[Wal82] Y. (Yehuda) Wallach. Alternating sequential/ parallel processing, volume 127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11194-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 329 pp. LCCN QA76.9.D5 .W19. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0127.htm;

http://www.springerlink.
com/content/978-0-387-
11194-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=127.

Wallis:1984:AST

[Wal84] Peter J. L. Wallis, editor.

Ada Software Tools Interfaces Workshop, Bath, July 13–15, 1983: proceedings, volume 180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1984. CO-DEN LNCSD9. ISBN 0-387-13878-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 A34 1983. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0180.htm;
http://www.springerlink.
com/content/978-0-387-
13878-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=180.

Walker:1999:CCI

[Wal99a] Michael Walker, editor.

Cryptography and Coding: 7th IMA Conference, Cirencester, UK, December 20–22, 1999: Proceedings, volume 1746 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66887-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1746. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1746.htm;
 http://www.springerlink.
com/content/978-3-540-
66887-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1746.

Walser:1999:IOL

[Wal99b] Joachim Paul Walser. Integer optimization by local search: a domainindependent approach, volume 1637 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66367-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 137 pp. LCCN T57.74 .W33 1999. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1637.
htm;   http://www.springerlink.
com/content/978-3-540-
66367-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1637.

Walsh:2001:PPC

[Wal01] Toby Walsh, editor.

Principles and Practice of Constraint Programming— CP 2001: 7th International Conference, CP 2001, Paphos, Cyprus, November 26–December 1, 2001. Proceedings, volume 2239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42863-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C7 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2239.htm;
http://www.springerlink.
com/content/978-3-540-
42863-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2239.

Wallace:2004:PPC

[Wal04a] M. (Mark) Wallace, editor. Principles and practice of constraint programming: CP 2004: 10th international conference, CP 2004, Toronto, Canada, September 27– October 1, 2004: proceedings, volume 3258 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23241-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.612 .C7 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23241-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3258;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100482.

Walukiewicz:2004:FSS

[Wal04b] Igor Walukiewicz, editor. Foundations of Software Science and Computation Structures: 7th International Conference, FOSSACS 2004, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2004, Barcelona, Spain, March 29–April 2, 2004: Proceedings, volume 2987 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21298-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F67 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2987.htm;

http://www.springerlink.
com/content/978-3-540-
21298-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2987;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95995.

Wansing:1993:LIS

[Wan93] Heinrich Wansing. The logic of information structures, volume 681 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 0-387-56734-8. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 163 pp. LCCN QA76 .W235 1993. US$35.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0681.
htm;   http://www.springerlink.
com/content/978-0-387-
56734-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=681.

Revision of the author’s doctoral thesis (Fachbereich Philosophie und Sozialwissenschaften I of the Free University of Berlin).

Wang:1998:ADG

[Wan98] Dongming Wang, editor. Automated deduction in geometry: International Workshop on Automated Deduction in Geometry, Toulouse, France, September 27– 29, 1996: Selected Papers, volume 1360 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64297-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA448.D38I577 1996.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1360.
htm;   http://www.springerlink.
com/content/978-3-540-
64297-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1360.

Wang:2001:CCA

[Wan01] Jie Wang, editor. Computing and combinatorics: 7th annual International Conference, COCOON 2001, Guilin, China, August 2001: Proceedings, volume 2108 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42494-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2108; QA75.5 .I5443 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2108.htm;
 http://www.springerlink.
com/content/978-3-540-
42494-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2108.

Wang:2003:RSF

[Wan03] Guoyin Wang, editor.

Rough sets, fuzzy sets, data mining, and granular computing: 9th International Conference, RSFDGrC 2003, Chongqing, China, May 26–29, 2003: Proceedings, volume 2639 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-14040-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .R494 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2639.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        14040-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2639.

Also available via the

World Wide Web.

Wang:2004:ATV

[Wan04] Farn Wang, editor. Automated Technology for Verification and Analysis: Second International Conference, ATVA 2004, Taipei, Taiwan, ROC, October 31–November 3, 2004, Proceedings, volume 3299 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23610-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3299.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23610-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3299;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b102065.

Wang:2005:FTN

[Wan05a] Farn Wang, editor. Formal Techniques for Networked and Distributed Systems — FORTE 2005: 25th IFIP WG 6.1 International Conference, Taipei, Taiwan, October 2–5, 2005. Proceedings, volume 3731 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29189-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29189-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3731.

Wang:2005:CCA

[Wan05b] Lusheng Wang, editor.

Computing and combinatorics: 11th annual international conference, COCOON 2005, Kunming, China, August 16– 29 [i.e., 16–19], 2005: proceedings, volume 3595 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28061-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .C549 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=3595;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        28061-3.

Woolf:2008:ITS

[WANL08] Beverley P. Woolf, Esma Aïmeur, Roger Nkambou, and Susanne Lajoie, editors. Intelligent Tutoring Systems: 9th International Conference, ITS 2008, Montreal, Canada, June 23– 27, 2008 Proceedings, volume 5091 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69130-8 (print), 3-540-69132-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        69132-7.

Warboys:1994:SPT

[War94a] Brian C. Warboys, editor. Software process technology: Third European workshop, EWSPT ’94, Villard de Lans, France, February 7– 9, 1994: proceedings, volume 772 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-57739-4 (New York), 3-540-57739-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .E987 1994. DM58.00. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0772.
htm;   http://www.springerlink.
com/content/978-0-387-
57739-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=772.

Wess:1994:TCB

[WAR94b] Stefan Wess, Klaus-Dieter Althoff, and Michael M. Richter, editors. Topics in case-based reasoning: first European workshop, EWCBR ’93, Kaiserslautern, Germany, November 1– 5, 1993: Selected Papers, volume 837 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58330-0 (Berlin), 0-387-58330-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E93 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0837.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        58330-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=837.

Wasniewski:1996:APC

[Waś96] Jerzy Waśniewski, editor. Applied parallel computing: industrialstrength computation and optimization: Third International Workshop, PARA 96, Lyngby, Denmark, August 18–21, 1996: proceedings, volume 1184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-62095-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1184.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62095-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1184.

Wade:2006:AHA

[WAS06] Vincent P. Wade, Helen Ashman, and Barry Smyth, editors. Adaptive Hypermedia and Adaptive Web-Based Systems: 4th International Conference, AH 2006, Dublin, Ireland, June 21–23, 2006. Proceedings, volume 4018 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34696-1 (print), 3-540-34697-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34697-5.

Watson:1995:PCB

[Wat95] Ian D. Watson, editor. Progress in casebased reasoning: first United Kingdom workshop, Salford, UK, January 12, 1995: proceedings, volume 1020 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60654-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 U52 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1020.htm;
 http://www.springerlink.
com/content/978-3-540-
60654-3;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1020.

Organized by AI-CBR and the British Computer Society Specialist Group on Expert Systems.

Watt:2000:CCI

[Wat00] David A. Watt, editor. Compiler construction: 9th International Conference, CC 2000, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2000, Berlin, Germany, March/April: Proceedings, volume 1781 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67263-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1781. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1781.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67263-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1781.

Wermter:2001:ENC

[WAW01] Stefan Wermter, James Austin, and David J.

Willshaw, editors. Emergent Neural Computational Architectures Based on Neuroscience: Towards Neuroscience-Inspired Computing, volume 2036 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42363-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 576 pp. LCCN QA267.A1 L43 no.2036; QA76.9.A73 E57 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2036.
htm;   http://www.springerlink.
com/content/978-3-540-
42363-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2036.

Wirsing:1989:AMT

[WB89] M. Wirsing and J. A.

Bergstra, editors. Algebraic methods: theory, tools, and applications, volume 394 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-51698-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 A44 1989. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0394.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        51698-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=394.

Proceedings of a workshop held on June 9– 11, 1987 in Passau, Germany.

Wangler:2000:AIS

[WB00] Benkt Wangler and Lars Bergman, editors. Advanced information systems engineering: 12th International Conference, CAiSE 2000, Stockholm, Sweden, June 5–9, 2000: Proceedings, volume 1789 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67630-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1789. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1789.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        67630-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=

0302-9743&volume=1789.

Wiederhold:2009:CIA

[WB09] Petra Wiederhold and Reneta P. Barneva, editors. Combinatorial Image Analysis: 13th International Workshop, IW-CIA 2009, Playa del Carmen, Mexico, November 24–27, 2009. Proceedings, volume 5852 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10208-5 (print), 3-642-10210-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10210-3.

Weyns:2008:EEM

[WBD08] Danny Weyns, Sven A.

Brueckner, and Yves Demazeau, editors. Engineering Environment-Mediated Multi-Agent Systems: International Workshop, EEMMAS 2007, Dresden, Germany, October 5, 2007. Selected Revised and Invited Papers, volume 5049 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85028-7 (print), 3-540-85029-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85029-8.

Wirsing:2008:SIS

[WBHR08] Martin Wirsing, Jean-Pierre Banâtre, Matthias Hölzl, and Axel Rauschmayer, editors. Software-Intensive Systems and New Computing Paradigms: Challenges and Visions, volume 5380 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89436-5 (print), 3-540-89437-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89437-7.

Wijers:1994:AIS

[WBW94] Gerard M. Wijers, Sjaak Brinkkemper, and Anthony I. Wasserman, editors. Advanced information systems engineering: 6th International Conference, CAiSE ’94, Utrecht, the Netherlands, June 6–10, 1994: proceedings, volume 811 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58113-8, 0-387-58113-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .N67 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0811.htm;
 http://www.springerlink.
com/content/978-3-540-
58113-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=811.

Wainer:1995:AAI

[WC95] Jacques Wainer and Ariadne Carvalho, editors. Advances in artificial intelligence: 12th Brazilian Symposium on Artificial Intelligence, SBIA ’95, Campinas, Brazil, October 10–12, 1995: proceedings, volume 991 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60436-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .S56 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0991.htm;
 http://www.springerlink.

                                        com/content/978-3-540-
                                        60436-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=991.

Wells:1998:MIC

[WC98] William M. Wells and A. C. F. (Alan C. F.) Colchester, editors. Medical image computing and computer-assisted intervention– MICCAI ’98: First International Conference, Cambridge, MA, USA, October 11– 13, 1998: Proceedings, volume 1496 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65136-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R857.O6I574 1998.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1496.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65136-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1496.

Woods:2008:RCA

[WCBD08] Roger Woods, Katherine Compton, Christos Bouganis, and Pedro C. Diniz, editors. Reconfigurable Computing: Architectures, Tools and Applications: 4th International Workshop, ARC 2008, London, UK, March 26–28, 2008. Proceedings, volume 4943 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78609-0 (print), 3-540-78610-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78610-8.

Wu:2005:ESS

[WCGB05] Zhaohui Wu, Chun Chen, Minyi Guo, and Jiajun Bu, editors. Embedded Software and Systems: First International Conference, ICESS 2004, Hangzhou, China, December 9–10, 2004, Revised Selected Papers, volume 3605 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28128-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
28128-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3605.

Wimmer:2010:EGI

[WCJS10] Maria A. Wimmer, Jean-Loup Chappelet, Marijn Janssen, and Hans J. Scholl, editors. Electronic Government: 9th IFIP WG 8.5 International Conference, EGOV 2010, Lausanne, Switzerland, August 29 — September 2, 2010. Proceedings, volume 6228 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14798-4 (print), 3-642-14799-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14799-9.

Whittle:2011:MDE

[WCK11] Jon Whittle, Tony Clark, and Thomas Kühne, editors. Model Driven Engineering Languages and Systems: 14th International Conference, MODELS 2011, Wellington, New Zealand, October 16–21, 2011. Proceedings, volume 6981 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24484-X (print), 3-642-24485-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24485-8.

Wang:2005:ANCa

[WCO05a] Lipo Wang, Ke Chen, and Yew S. Ong, editors. Advances in natural computation: first international conference, ICNC 2005, Changsha, China, August 27–29, 2005: proceedings, volume 3610 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28323-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I26 2005. URL http:

//www.springerlink.
com/content/978-3-540-
28323-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3610.

Wang:2005:ANCb

[WCO05b] Lipo Wang, Ke Chen, and Yew S. Ong, editors. Advances in natural computation: first international conference, ICNC 2005, Changsha, China, August 27–29, 2005: proceedings, volume 3611 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28325-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I26 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28325-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3611.

Wang:2005:ANCc

[WCO05c] Lipo Wang, Ke Chen, and Yew S. Ong, editors. Advances in natural computation: first international conference, ICNC 2005, Changsha, China, August 27–29, 2005: proceedings, volume 3612 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28320-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I26 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        28320-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3612.

Wyatt:2000:ARL

[WD00] Jeremy Wyatt and John Demiris, editors. Advances in robot learning: 8th European Workshop on Learning Robots, EWLR-8, Lausanne, Switzerland, September 18, 1999: Proceedings, volume 1812 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41162-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1812. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1812.htm;
 http://www.springerlink.
com/content/978-3-540-
41162-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1812.

Wu:2010:COAa

[WD10a] Weili Wu and Ovidiu Daescu, editors. Combinatorial Optimization and Applications: 4th International Conference, COCOA 2010, Kailua-Kona, HI, USA, December 18–20, 2010, Proceedings, Part I, volume 6508 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17457-4 (print), 3-642-17458-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17458-2.

Wu:2010:COAb

[WD10b] Weili Wu and Ovidiu Daescu, editors. Combinatorial Optimization and Applications: 4th International Conference, COCOA 2010, Kailua-Kona, HI, USA, December 18–20, 2010, Proceedings, Part II, volume 6509 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17460-4 (print), 3-642-17461-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        17461-2.

Wang:2010:AICa

[WDGL10a] Fu Lee Wang, Hepu Deng, Yang Gao, and Jingsheng Lei, editors. Artificial Intelligence and Computational Intelligence: International Conference, AICI 2010, Sanya, China, October 23–24, 2010, Proceedings, Part I, volume 6319 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16529-X (print), 3-642-16530-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16530-6.

Wang:2010:AICb

[WDGL10b] Fu Lee Wang, Hepu Deng, Yang Gao, and Jingsheng Lei, editors. Artificial Intelligence and Computational Intelligence: International Conference, AICI 2010, Sanya, China, October 23–24, 2010, Proceedings, Part II, volume 6320 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16526-5 (print), 3-642-16527-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16527-6.

Winter:2007:SIT

[WDKK07] Stephan Winter, Matt Duckham, Lars Kulik, and Ben Kuipers, editors. Spatial Information Theory: 8th International Conference, COSIT 2007, Melbourne, Australiia, September 19–23, 2007. Proceedings, volume 4736 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74786-9 (print), 3-540-74788-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74788-8.

Wyrzykowski:2008:PPA

[WDKW08] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Wasniewski, editors. Parallel Processing and Applied Mathematics: 7th International Conference, PPAM 2007, Gdansk, Poland, September 9– 12, 2007 Revised Selected Papers, volume 4967 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68105-1 (print), 3-540-68111-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68111-3.

Wyrzykowski:2010:PPAa

[WDKW10a] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Wasniewski, editors. Parallel Processing and Applied Mathematics: 8th International Conference, PPAM 2009, Wroclaw, Poland, September 13– 16, 2009. Revised Selected Papers, Part I, volume 6067 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14389-X (print), 3-642-14390-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14390-8.

Wyrzykowski:2010:PPAb

[WDKW10b] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Wasniewski, editors. Parallel Processing and Applied Mathematics: 8th International Conference, PPAM 2009, Wroclaw, Poland, September 13– 16, 2009, Revised Selected Papers, Part II, volume 6068 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14402-0 (print), 3-642-14403-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14403-5.

Wyrzykowski:2012:PPAa

[WDKW12a] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Waśniewski, editors. Parallel Processing and Applied Mathematics: 9th International Conference, PPAM 2011, Torun, Poland, September 11– 14, 2011. Revised Selected Papers, Part I, volume 7203 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31463-5 (print), 3-642-31464-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31464-3.

Wyrzykowski:2012:PPAb

[WDKW12b] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Waśniewski, editors. Parallel Processing and Applied Mathematics: 9th International Conference, PPAM 2011, Torun, Poland, September 11– 14, 2011. Revised Selected Papers, Part II, volume 7204 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31499-6 (print), 3-642-31500-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31500-8.

Wyrzykowski:2006:PPA

[WDM06] Roman Wyrzykowski, Jack Dongarra, and Norbert Meyer, editors. Parallel Processing and Applied Mathematics: 6th International Conference, PPAM 2005, Poznan, Poland, September 11–14, 2005, Revised Selected Papers, volume 3911 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34141-2 (soft cover). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34141-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3911.

Wyrzykowski:2002:PPA

[WDPW02] Roman Wyrzykowski, Jack Dongarra, Marcin Paprzycki, and Jerzy Wasńiewski, editors. Parallel Processing and Applied Mathematics: 4th International Conference, PPAM 2001 Naleczów, Poland, September 9–12, 2001. Revised Papers, volume 2328 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43792-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P69 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2328.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43792-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2328.

Wyrzykowski:2004:PPA

[WDPW04] Roman Wyrzykowski, Jack Dongarra, Marcin Paprzycki, and Jerzy Waśniewski, editors. Parallel Processing and Applied Mathematics: 5th International Conference, PPAM 2003, Czestochowa, Poland, September 7–10, 2003: Revised Papers, volume 3019 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21946-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3019.htm;
 http://www.springerlink.
com/content/978-3-540-
21946-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3019;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97218.

Weihrauch:1979:TCS

[Wei79] K. Weihrauch, editor.

Theoretical computer science: 4th GI conference, Aachen, March 26– 28, 1979, volume 67 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1979. CODEN LNCSD9. ISBN 0-387-09118-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .G181 1979. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0067.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        09118-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=67.

English, French, or Ger-

man.

Weiss:1997:DAI

[Wei97] Gerhard Weiss, editor. Distributed artificial intelligence meets machine learning: learning in multi-agent environments: ECAI’96 Workshop LDAIS, Budapest, Hungary, August 13, 1996, ICMAS’96 Workshop LIOME, Kyoto, Japan, December 10, 1996, Selected Papers, volume 1221 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62934-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .D57 1997. URL http:

                                        //link.springer-ny.

com/link/service/series/
0558/tocs/t1221.htm;
http://www.springerlink.
com/content/978-3-540-
62934-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1221.

Welch:2001:DCI

[Wel01] Jennifer Welch, editor.

Distributed computing: 15th International Conference, DISC 2001, Lisbon, Portugal, October 3–5, 2001: Proceedings, volume 2180 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42605-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I597 2001; QA267.A1 L43 no.2180. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2180.htm;
http://www.springerlink.
com/content/978-3-540-
42605-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2180.

Westfechtel:1999:MTM

[Wes99] Bernhard Westfechtel.

Models and Tools for Managing Development Processes, volume 1646 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66756-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 418 pp. LCCN QA76.76.D47 W475 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1646.
htm;   http://www.springerlink.
com/content/978-3-540-
66756-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1646.

Westort:2001:DEM

[Wes01] Caroline Y. Westort, editor. Digital Earth Moving: First International Symposium, DEM 2001, Manno, Switzerland, September 5–7, 2001: Proceedings, volume 2181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42586-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA715 .D54 2001; QA267.A1 L43 no.2181. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2181.htm;
 http://www.springerlink.

com/content/978-3-540-
42586-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2181.

Wetter:1992:CDK

[Wet92] Thomas Wetter, editor.

Current developments in knowledge acquisition: EKAW’92: 6th European Knowledge Acquisition Workshop, Heidelberg and Kaiserslautern, Germany, May 18–22, 1992: proceedings, volume 599 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55546-3 (Berlin), 0-387-55546-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E36 1992. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0599.htm;
http://www.springerlink.
com/content/978-3-540-
55546-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=599.

Williams:1977:DIP

[WF77] John H. Williams and David A. Fisher, editors. Design and implementation of programming languages: proceedings of a DoD sponsored workshop, October, 1976, Ithaca, volume 54 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CODEN LNCSD9. ISBN 0-387-08360-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .D47. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0054.htm;
 http://www.springerlink.
com/content/978-0-387-
08360-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=54.

Wachsmuth:1998:GSL

[WF98] Ipke Wachsmuth and Martin Fröhlich, editors. Gesture and sign language in humancomputer interaction: International Gesture Workshop, Bielefeld, Germany, September 17– 19, 11997: Proceedings, volume 1371 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64424-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.9.H85I618 1997. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1371.htm;
http://www.springerlink.
com/content/978-3-540-
64424-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1371.

Winckler:2012:HCS

[WFB12] Marco Winckler, Peter Forbrig, and Regina Bernhaupt, editors. Human-Centered Software Engineering: 4th International Conference, HCSE 2012, Toulouse, France, October 29–31, 2012. Proceedings, volume 7623 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34346-5 (print), 3-642-34347-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34347-6.

Wang:2009:HLE

[WFZL09] Fu Lee Wang, Joseph Fong, Liming Zhang, and Victor S. K. Lee, editors. Hybrid Learning and Education: Second International Conference, ICHL 2009, Macau, China, August 25–27, 2009. Proceedings, volume 5685 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03696-1 (print), 3-642-03697-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03697-2.

Weyns:2011:AOS

[WG11] Danny Weyns and Marie-Pierre Gleizes, editors. Agent-Oriented Software Engineering XI: 11th International Workshop, AOSE 2010, Toronto, Canada, May 10–11, 2010, Revised Selected Papers, volume 6788 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22635-3 (print), 3-642-22636-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22636-6.

Winther:2005:CSR

[WGD05] Rune Winther, Bjørn Axel Gran, and Gustav Dahll, editors. Computer safety, reliability, and security: 24th international conference, SAFE-COMP 2005, Fredrikstad, Norway, September 28–30, 2005: proceedings, volume 3688 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29200-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ212.2 .I545 2005; QA76.76.R44 S23 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29200-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3688.

Wierse:1996:DID

[WGL96] Andreas Wierse, Georges G.

Grinstein, and Ulrich Lang, editors. Database issues for data visualization: IEEE Visualization ’95 Workshop, Atlanta, Georgia, USA, October 28, 1995: proceedings, volume 1183 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-62221-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I15 1995. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1183.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        62221-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1183.

Wang:2010:WIS

[WGLL10] Fu Lee Wang, Zhiguo Gong, Xiangfeng Luo, and Jingsheng Lei, editors. Web Information Systems and Mining: International Conference, WISM 2010, Sanya, China, October 23–24, 2010. Proceedings, volume 6318 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16514-1 (print), 3-642-16515-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16515-3.

Wang:2009:TSD

[WGMP09] Qing Wang, Vahid Garousi, Raymond Madachy, and Dietmar Pfahl, editors. Trustworthy Software Development Processes: International Conference on Software Process, ICSP 2009 Vancouver, Canada, May 16–17, 2009 Proceedings, volume 5543 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01679-0 (print), 3-642-01680-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01680-6.

Wiedermann:1995:MFC

[WH95] J. Wiedermann and Petr Hajek, editors. Mathematical foundations of computer science 1995: 20th International Symposium, MFCS ’95, Prague, Czech Republic, August 28–September 1, 1995: proceedings, volume 969 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60246-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 I54 1995. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0969.htm;
 http://www.springerlink.
com/content/978-3-540-
60246-0;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=969.

Wirsing:2011:RSE

[WH11] Martin Wirsing and Matthias Hölzl, editors. Rigorous Software Engineering for Service-Oriented Systems: Results of the SENSO-RIA Project on Software Engineering for Service-Oriented Computing, volume 6582 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20400-7 (print), 3-642-20401-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20401-2.

Whang:2003:AKD

[Wha03] Kyu-Young Whang, editor. Advances in knowledge discovery and data mining: 7th Pacific-Asia Conference, PAKDD 2003, Seoul. Korea, April 30–May 2, 2003: Proceedings, volume 2637 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-04760-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 P17 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2637.htm;
 http://www.springerlink.
com/content/978-3-540-
04760-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2637.

Also available via the World Wide Web.

Weske:2007:WIS

[WHG07] Mathias Weske, Mohand-Saïd Hacid, and Claude Godart, editors. Web Information Systems Engineering — WISE 2007 Workshops: WISE 2007 International Workshops Nancy, France, December 3, 2007 Proceedings, volume 4832 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77009-7 (print), 3-540-77010-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77010-7.

Whitesides:1998:GDI

[Whi98] Sue H. Whitesides, editor. Graph Drawing: 6th International Symposium, GD’98, Montréal, Canada, August 13– 15, 1998. Proceedings, volume 1547 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65473-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 1998.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1547.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        65473-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1547.

Proceedings of the 6th Symposium on Graph Drawing (GD’98).

White:2000:EMT

[Whi00] John S. White, editor. Envisioning Machine Translation in the Information Future: 4th Conference of the Association for Machine Translation in the Americas, AMTA 2000, Cuernavaca, Mexico, October 10–14, 2000. Proceedings, volume 1934 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41117-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1934.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1934.
htm;   http://www.springerlink.
com/content/978-3-540-
41117-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1934.

Wada:2009:AIV

[WHL09] Toshikazu Wada, Fay Huang, and Stephen Lin, editors. Advances in Image and Video Technology: Third Pacific Rim Symposium, PSIVT 2009, Tokyo, Japan, January 13–16, 2009. Proceedings, volume 5414 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-540-92956-8 (print), 3-540-92957-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
92957-4.

Wirsing:2010:TGC

[WHR10] Martin Wirsing, Martin Hofmann, and Axel Rauschmayer, editors. Trustworthly Global Computing: 5th International Symposium, TGC 2010, Munich, Germany, February 24–26, 2010, Revised Selected Papers, volume 6084 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15639-8 (print), 3-642-15640-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        15640-3.

Wolf:2001:QSI

[WHS01] Lars Christian Wolf, David Hutchison, and Ralf Steinmetz, editors. Quality of service — IWQoS 2001: 9th International Workshop, Karlsruhe, Germany, June 6–8, 2001: Proceedings, volume 2092 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42217-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Electronic book; QA267.A1 L43 no.2092; TK5105.5 .I97 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2092.htm;
 http://www.springerlink.
com/content/978-3-540-
42217-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2092.

Wiener:1999:ACC

[Wie99] Michael Wiener, editor. Advances in cryptology — CRYPTO ’99: 19th Annual International Cryptology Conference, Santa Barbara, California, USA, August 15–19, 1999 proceedings, volume 1666 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66347-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1666.
htm;   http://www.springerlink.
com/content/978-3-540-
66347-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1666.

Wiedijk:2006:SPW

[Wie06] Freek Wiedijk, editor. The Seventeen Provers of the World, volume 3600 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30704-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 157 pp. LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        30704-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3600.

Foreword by Dana S.

Scott.

Wiil:2004:CMM

[Wii04] Uffe Kock Wiil, editor. Computer Music Modeling and Retrieval: International Symposium, CMMR 2003, Montpellier, France, May 26– 27, 2003: Revised Papers, volume 2771 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-20922-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ML1380 .C645 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2771.htm;

http://www.springerlink.
com/content/978-3-540-
20922-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2771;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12000.

Wiil:2005:CMM

[Wii05a] Uffe Kock Wiil, editor.

Computer music modeling and retrieval: Second International Symposium, CMMR 2004, Esbjerg, Denmark, May 26–29, 2004: revised papers, volume 3310 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24458-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ML74 .C+. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3310;
http://www.springerlink.
com/content/978-3-540-
24458-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3310;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b105507.

Wiil:2005:MIS

[Wii05b] Uffe Kock Wiil, editor. Metainformatics: International Symposium, MIS 2004, Salzburg, Austria, September 15–18, 2004, Revised Selected Papers, volume 3511 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27328-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .M55 2004. URL http:

//www.springerlink.
com/content/978-3-540-
27328-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3511.

Williams:1986:ACC

[Wil86] Hugh C. Williams, editor. Advances in cryptology — CRYPTO ’85: proceedings: August 18– 22, 1985, volume 218 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16463-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C791 1985; QA267.A1 L43 no.218. URL http:

//link.springer-ny.
com/link/service/series/

0558/tocs/t0218.htm;
http://www.springerlink.
com/content/978-0-387-
16463-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=218.

Williams:1991:ITF

[Wil91] James G. Williams. Instantiation theory: on the foundations of automated deduction, volume 518 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54333-3 (Berlin), 0-387-54333-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 133 pp. LCCN QA76.9.A96 W55 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0518.
htm;   http://www.springerlink.
com/content/978-3-540-
54333-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=518.

Williams:1999:QCQ

[Wil99] Colin P. Williams, editor. Quantum computing and quantum communications: First NASA International Conference, QCQS [i.e., QCQC] ’98, Palm Springs, California, USA, February 17– 20, 1998: Selected Papers, volume 1509 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65514-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.889 .Q37 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1509.htm;
 http://www.springerlink.
com/content/978-3-540-
65514-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1509.

Wilhelm:2001:CCI

[Wil01a] R. Wilhelm, editor. Compiler construction: 10th International Conference, CC 2001, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2001, Genova, Italy, April 2– 6, 2001: Proceedings, volume 2027 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41861-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2027.
htm;   http://www.springerlink.
com/content/978-3-540-
41861-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2027.

Wilhelm:2001:IYB

[Wil01b] Reinhold Wilhelm. Informatics 10 years back, 10 years ahead, volume 2000 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41635-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 368 pp. LCCN QA76 .I28 2001; QA267.A1 L43 no.2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2000.htm;
http://www.springerlink.
com/content/978-3-540-
41635-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2000.

Wimmer:2003:KME

[Wim03] Maria A. Wimmer, editor. Knowledge Management in Electronic Government: 4th IFIP International Working Conference, KMGov 2003, Rhodes, Greece, May 26–28, 2003: Proceedings, volume 2645 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40145-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN JF1525.A8 K64 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2645.htm;
 http://www.springerlink.
com/content/978-3-540-
40145-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2645.

Wimmer:2004:KME

[Wim04] Maria A. Wimmer, editor. Knowledge Management in Electronic Government: 5th IFIP International Working Conference, KMGov 2004, Krems, Austria, May 17–19, 2004: Proceedings, volume 3035 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22002-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t3035.htm;
http://www.springerlink.
com/content/978-3-540-
22002-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3035;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b97726.

Winkowski:1978:MFC

[Win78] Józef Winkowski, editor. Mathematical foundations of computer science, 1978: proceedings, 7th Symposium, Zakopane, Poland, September 4–8, 1978, volume 64 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1978. CODEN LNCSD9. ISBN 0-387-08917-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S9194 1978. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0064.htm;
http://www.springerlink.
com/content/978-0-387-
08917-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=64.

Winkler:2004:ADG

[Win04] Franz Winkler, editor.

Automated Deduction in Geometry: 4th International Workshop, ADG 2002, Hagenberg Castle, Austria, September 4– 6, 2002: Revised Papers, volume 2930 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20927-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA448.D38 I577 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2930.
htm;   http://www.springerlink.
com/content/978-3-540-
20927-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2930;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95516.

Winslett:2009:SSD

[Win09] Marianne Winslett, editor. Scientific and Statistical Database Management: 21st International Conference, SS-DBM 2009 New Orleans, LA, USA, June 2–4, 2009 Proceedings, volume 5566 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02278-2 (print), 3-642-02279-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02279-1.

Winkler:2011:AII

[Win11] Franz Winkler, editor. Algebraic Informatics: 4th International Conference, CAI 2011, Linz, Austria, June 21– 24, 2011. Proceedings, volume 6742 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21492-4 (print), 3-642-21493-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21493-6.

Wooldridge:1995:IAE

[WJ95] Michael J. Wooldridge and Nick Jennings, editors. Intelligent agents: ECAI-94 Workshop on Agent Theories, Architectures, and Languages, Amsterdam, the Netherlands, August 8–9, 1994: proceedings, volume 890 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-58855-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q336 .E23 1994. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0890.
htm;   http://www.springerlink.
com/content/978-3-540-
58855-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=890.

Wang:2005:FSK

[WJ05] Lipo Wang and Yaochu Jin, editors. Fuzzy Systems and Knowledge Discovery: Second International Conference, FSKD 2005, Changsha, China, August 27–29, 2005: Proceedings, volume 3613–3614 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28312-9 (pt. 1), 3-540-28331-5 (pt. 2). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA248 .F797 2005. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3613;
 http://springerlink.

metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3614;
http://www.springerlink.
com/content/978-3-540-
28312-6.

Wang:2012:WAI

[WJL^{+}12] Liwei Wang, Jingjue Jiang, Jiaheng Lu, Liang Hong, and Bin Liu, editors. Web-Age Information Management: WAIM 2011 International Workshops: WGIM 2011, XMLDM 2011, SNA 2011, Wuhan, China, September 14–16, 2011, Revised Selected Papers, volume 7142 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28634-8 (print), 3-642-28635-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
28635-3.

Winckler:2007:TMD

[WJP07] Marco Winckler, Hilary Johnson, and Philippe Palanque, editors. Task Models and Diagrams for User Interface Design: 6th International Workshop, TAMODIA 2007, Toulouse, France, November 7–9, 2007.

Proceedings, volume 4849 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77221-9 (print), 3-540-77222-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77222-4.

Wang:2006:FSK

[WJS^{+}06] Lipo Wang, Licheng Jiao, Guanming Shi, Xue Li, and Jing Liu, editors. Fuzzy Systems and Knowledge Discovery: Third International Conference, FSKD 2006, Xi’an, China, September 24–28, 2006. Proceedings, volume 4223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-45916-2 (print), 3-540-45917-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
45917-0.

Weber:2002:FAS

[WK02] Herbert Weber and Ralf-Detlef Kutsche, editors. Fundamental approaches to software engineering: 5th International Conference, FASE 2002, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2002, Grenoble, France, April 8–12, 2002: Proceedings, volume 2306 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43353-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F44 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2306.htm;
 http://www.springerlink.
com/content/978-3-540-
43353-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2306.

Won:2006:ISC

[WK06] Dongho Won and Seungjoo Kim, editors. Information Security and Cryptology — ICISC 2005: 8th International Conference, Seoul, Korea, December 1–2, 2005, Revised Selected Papers, volume 3935 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33354-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        33354-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3935.

Wachsmuth:2008:MCR

[WK08] Ipke Wachsmuth and Günther Knoblich, editors. Modeling Communication with Robots and Virtual Humans: Second ZiF Research Group International Workshop on Embodied Communication in Humans and Machines, Bielefeld, Germany, April 5–8, 2006, Revised Selected Papers, volume 4930 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79036-5 (print), 3-540-79037-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79037-2.

Wirsing:2004:RIS

[WKB04] Martin Wirsing, Alexander Knapp, and Simonetta Balsamo, editors. Radical Innovations of Software and Systems Engineering in the Future: 9th International Workshop, RISSEF 2002, Venice, Italy, October 7–11, 2002: Revised Papers, volume 2941 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21179-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .R49 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2941.htm;
http://www.springerlink.
com/content/978-3-540-
21179-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2941;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96009.

Wyeld:2008:VSM

[WKD08] Theodor G. Wyeld, Sarah Kenderdine, and Michael Docherty, editors. Virtual Systems and Multimedia: 13th International Conference, VSMM 2007, Brisbane, Australia, September 23–26, 2007, Revised Selected Papers, volume 4820 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78565-5 (print), 3-540-78566-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78566-8.

Wu:1998:RDK

[WKK98] Xindong Wu, Ramamohanarao Kotagiri, and Kevin B. Korb, editors. Research and development in knowledge discovery and data mining: Second Pacific-Asia Conference, PAKDD-98, Melbourne, Australia, April 15–17, 1998: Proceedings, volume 1394 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64383-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3P17 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1394.htm;
 http://www.springerlink.
com/content/978-3-540-
64383-8;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1394.

Walter:2003:CHE

[WKP03] Colin D. Walter, Çetin K.

Koç, and Christof Paar, editors. Cryptographic Hardware and Embedded Systems—CHES 2003: 5th International Workshop, Cologne, Germany, September 8–10, 2003: Proceedings, volume 2779 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40833-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t2779.
htm;   http://www.springerlink.
com/content/978-3-540-
40833-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2779;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13240.

Wolpers:2010:STI

[WKS^{+}10] Martin Wolpers, Paul A.

Kirschner, Maren Scheffel, Stefanie Lindstaedt, and Vania Dimitrova, editors. Sustaining TEL: From Innovation to Learning and Practice: 5th European Conference on Technology Enhanced Learning, EC-TEL 2010, Barcelona, Spain, September 28 — October 1, 2010. Proceedings, volume 6383 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16019-0 (print), 3-642-16020-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16020-2.

Woodcock:1993:FIS

[WL93] Jim Woodcock and P. G. Larsen, editors. FME ’93: industrial-strength formal methods: First International Symposium of Formal Methods Europe, Odense, Denmark, April 19–23, 1993 proceedings, volume 670 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-56662-7 (Berlin), 0-387-56662-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I593 1993. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t0670.htm;
http://www.springerlink.
com/content/978-3-540-
56662-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=670.

Winkler:2003:SNS

[WL03] Franz Winkler and Ulrich Langer, editors. Symbolic and numerical scientific computation: Second International Conference, SNSC 2001, Hagenberg, Austria, September 12– 14, 2001: Revised Papers, volume 2630 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40554-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA370 .S67 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2630.htm;
http://www.springerlink.
com/content/978-3-540-
40554-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2630.

Weske:2004:OOI

[WL04] Mathias Weske and Peter Liggesmeyer, editors. Object-Oriented and Internet-Based Technologies: 5th Annual International Conference on Object-Oriented and Internet-Based Technologies, Concepts, and Applications for a Networked World, Net.ObjectDays 2004, Erfurt, Germany, September 27–30, 2004: Proceedings, volume 3263 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23201-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .N632 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23201-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3263.

Wang:2006:SEL

[WLC^{+}06] Tzai-Der Wang, Xiaodong Li, Shu-Heng Chen, Xufa Wang, Hussein Abbass, Hitoshi Iba, Guo-Liang Chen, and Xin Yao, editors. Simulated Evolution and Learning: 6th International Conference, SEAL 2006, Hefei, China, October 15–18, 2006. Proceedings, volume 4247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47331-9 (print), 3-540-47332-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
47332-9.

Wang:2008:RSK

[WLGB^{+}08] Guoyin Wang, Tianrui Li, Jerzy W. Grzymala-Busse, Duoqian Miao, Andrzej Skowron, and Yiyu Yao, editors. Rough Sets and Knowledge Technology: Third International Conference, RSKT 2008, Chengdu, China, May 17–19, 2008. Proceedings, volume 5009 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79720-3 (print), 3-540-79721-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
79721-0.

Lee:1999:DAS

[wLN99] S.-W. (Seong whan) Lee and Yasuaki Nakano, editors. Document analysis systems: theory and practice: Third IAPR Workshop, DAS’98, Nagano, Japan, November 4–6, 1998: Selected Papers, volume 1655 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66507-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1640 .D37 1998. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1655.htm;
 http://www.springerlink.
com/content/978-3-540-
66507-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1655.

Wang:2011:WAI

[WLO^{+}11] Haixun Wang, Shijun Li, Satoshi Oyama, Xiaohua Hu, and Tieyun Qian, editors. Web-Age Information Management: 12th International Conference, WAIM 2011, Wuhan, China, September 14–16, 2011. Proceedings, volume 6897 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23534-4 (print), 3-642-23535-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23535-1.

Wen:2009:RSK

[WLP^{+}09] Peng Wen, Yuefeng Li, Lech Polkowski, Yiyu Yao, Shusaku Tsumoto, and Guoyin Wang, editors. Rough Sets and Knowledge Technology: 4th International Conference, RSKT 2009, Gold Coast, Australia, July 14–16, 2009. Proceedings, volume 5589 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02961-2 (print), 3-642-02962-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02962-2.

Wu:2008:ISI

[WLRL08] Tzong-Chen Wu, Chin-Laung Lei, Vincent Rijmen, and Der-Tsai Lee, editors. Information Security: 11th International Conference, ISC 2008, Taipei, Taiwan, September 15–18, 2008. Proceedings, volume 5222 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85884-9 (print), 3-540-85886-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85886-7.

Wolf:1989:PFI

[WLS89] Gottfried Wolf, Tamas Legendi, and Udo Schendel, editors. Parcella ’88: Fourth International Workshop on Parallel Processing by Cellular Automata and Arrays, Berlin, GDR, October 17–21, 1988: proceedings, volume 342 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CO-DEN LNCSD9. ISBN 0-387-50647-0 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.5 .I621 1988. DM50.00. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t0342.
                                        htm;   http://www.springerlink.
                                        com/content/978-0-387-
                                        50647-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=342.

Wang:2005:ANNa

[WLY05a] Jun Wang, Xiaofeng Liao, and Zhang Yi, editors. Advances in Neural Networks: ISNN 2005: Second International Symposium on Neural Networks, Chongqing, China, May 30–June 1, 2005, Proceedings, Part I, volume 3496 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25912-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
25912-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3496;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b136476.

Wang:2005:ANNb

[WLY05b] Jun Wang, Xiaofeng Liao, and Zhang Yi, editors. Advances in Neural Networks: ISNN 2005: Second International Symposium on Neural Networks Chongqing, China, May 30–June 1, 2005 Proceedings, Part II, volume 3497 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25913-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25913-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3497;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136478.

Wang:2005:ANNc

[WLY05c] Jun Wang, Xiaofeng Liao, and Zhang Yi, editors. Advances in Neural Networks: ISNN 2005: Second International Symposium on Neural Networks Chongqing, China, May 30–June 1, 2005 Proceedings, Part III, volume 3498 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25914-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25914-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3498;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b136479.

Wilson:2003:MSI

[WM03] Michael J. Wilson and Ralph R. Martin, editors. Mathematics of Surfaces: 10th IMA International Conference, Leeds, UK, September 15–17, 2003: Proceedings, volume 2768 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20053-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA571 .I43 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2768.htm;
http://www.springerlink.
com/content/978-3-540-
20053-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2768;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11966.

Williams:2004:KTD

[WM04] Howard Williams and Lachlan MacKinnon, editors. Key Technologies for Data Management: 21st British National Conference on Databases, BNCOD 21, Edinburgh, UK, July 7– 9, 2004: Proceedings, volume 3112 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22382-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 B75 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3112.
htm;   http://www.springerlink.
com/content/978-3-540-
22382-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3112;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98625.

Wong:2010:NIPb

[WMB10a] Kok Wai Wong, B. Sumudu U. Mendis, and Abdesselam Bouzerdoum, editors. Neural Information Processing. Models and Applications: 17th International Conference, ICONIP 2010, Sydney, Australia, November 22– 25, 2010, Proceedings, Part II, volume 6444 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17533-3 (print), 3-642-17534-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17534-3.

Wong:2010:NIPa

[WMB10b] Kok Wai Wong, B. Sumudu U.

Mendis, and Abdesselam Bouzerdoum, editors. Neural Information Processing. Theory and Algorithms: 17th International Conference, ICONIP 2010, Sydney, Australia, November 22– 25, 2010, Proceedings, Part I, volume 6443 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-17536-8 (print), 3-642-17537-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
17537-4.

Wang:2007:CIS

[WmCL07] Yuping Wang, Yiu ming Cheung, and Hailin Liu, editors. Computational Intelligence and Security: International Conference, CIS 2006. Guangzhou, China, November 3–6, 2006. Revised Selected Papers, volume 4456 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74376-6 (print), 3-540-74377-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74377-4.

Weyns:2010:SOA

[WMdLA10] Danny Weyns, Sam Malek, Rogério de Lemos, and Jesper Andersson, editors. Self-Organizing Architectures: First International Workshop, SOAR 2009, Cambridge, UK, September 14, 2009, Revised Selected and Invited Papers, volume 6090 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14411-X (print), 3-642-14412-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14412-7.

Welzl:2000:ALP

[WMR00] Emo Welzl, Ugo Montanari, and Jose D. P. Rolim, editors. Automata, languages and programming: 27th International Colloquium, ICALP 2000, Geneva, Switzerland, July 9– 15, 2000: Proceedings, volume 1853 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CODEN LNCSD9. ISBN 3-540-67715-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1853.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1853.
htm;   http://www.springerlink.
com/content/978-3-540-
67715-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1853.

Wermelinger:2004:FAS

[WMS04] Michel Wermelinger and Tiziana Margaria-Steffen, editors. Fundamental Approaches to Software Engineering: 7th International Conference, FASE 2004, held as part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2004, Barcelona, Spain, March 29–April 2, 2004: Proceedings, volume 2984 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21305-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F44 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2984.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        21305-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2984;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b95935.

Wooldridge:1996:IAI

[WMT96] Michael J. Wooldridge, Jorg P. Muller, and Milind Tambe, editors. Intelligent agents II: agent theories, architectures, and languages: IJCAI ’95 workshop (ATAL), Montréal, Canada, August 19– 20, 1995: proceedings, volume 1037 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60805-2 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q336 .I67 1996. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1037.
                                        htm;   http://www.springerlink.

com/content/978-3-540-
60805-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1037.

Wirsing:1996:AMS

[WN96] M. (Martin) Wirsing and M. Nivat, editors. Algebraic methodology and software technology: 5th International Conference, AMAST ’96, Munich, Germany, July 1– 5, 1996: proceedings, volume 1101 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61463-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I5713 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1101.htm;
http://www.springerlink.
com/content/978-3-540-
61463-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1101.

Widmayer:1999:GTC

[WNE99] Peter Widmayer, Gabriele Neyer, and Stephan Eidenbenz, editors. Graphtheoretic concepts in computer science: 25th International Workshop, WG’99, Ascona, Switzerland, June 17–19, 1999:

Proceedings, volume 1665 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66731-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1665.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1665.
htm;   http://www.springerlink.
com/content/978-3-540-
66731-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1665.

Winkler:2005:DSM

[WNL05] Joab Winkler, Mahesan Niranjan, and Neil Lawrence, editors. Deterministic and Statistical Methods in Machine Learning: First International Workshop, Sheffield, UK, September 7–10, 2004. Revised Lectures, volume 3635 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29073-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29073-5;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3635.

Wolper:1995:CAV

[Wol95] Pierre Wolper, editor.

Computer aided verification: 7th International Conference, CAV ’95, Liege, Belgium, July 3–5, 1995: proceedings, volume 939 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60045-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 C38 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0939.
htm;   http://www.springerlink.
com/content/978-3-540-
60045-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=939.

Wolter:2007:FMS

[Wol07] Katinka Wolter, editor. Formal Methods and Stochastic Models for Performance Evaluation: Fourth European Performance Engineering Workshop, EPEW 2007, Berlin, Germany, September 27–28, 2007. Proceedings, volume 4748 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75210-2 (print), 3-540-75211-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75211-0.

Won:2001:ISC

[Won01] Dongho Won, editor. Information security and cryptology: ICISC 2000, Third International Conference, Seoul, Korea, December 8–9, 2000: Proceedings, volume 2015 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-41782-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I32 2000; QA267.A1 L43 no.2015. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2015.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41782-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2015.

Wood:1980:GFI

[Woo80] Derick Wood. Grammar and L forms: an introduction, volume 91 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1980. CODEN LNCSD9. ISBN 0-387-10233-7 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 314 pp. LCCN QA267.3 .W66. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0091.htm;
 http://www.springerlink.
com/content/978-0-387-
10233-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=91.

Weichselberger:1990:MUK

[WP90] Kurt Weichselberger and Sigrid Pohlmann. A methodology for uncertainty in knowledge-based systems, volume 419 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CO-DEN LNCSD9. ISBN 3-540-52336-7 (Berlin), 0-387-52336-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). 132 pp. LCCN QA76.76.E95 W44 1990. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0419.htm;
 http://www.springerlink.
com/content/978-3-540-

                                        52336-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=419.

Wieringa:2010:REF

[WP10] Roel Wieringa and Anne Persson, editors. Requirements Engineering: Foundation for Software Quality: 16th International Working Conference, REFSQ 2010, Essen, Germany, June 30– July 2, 2010. Proceedings, volume 6182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14191-9 (print), 3-642-14192-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14192-8.

Wolff:2004:CSW

[WPD04] Karl Erich Wolff, Heather D. Pfeiffer, and Harry S. Delugach, editors. Conceptual Structures at Work: 12th International Conference on Conceptual Structures, ICCS 2004, Huntsville, AL, USA, July 19– 23, 2004: Proceedings, volume 3127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22392-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387.2 .I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3127.htm;
 http://www.springerlink.
com/content/978-3-540-
22392-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3127;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98793.

Wermter:2005:BNL

[WPE05] Stefan Wermter, Günther Palm, and Mark Elshaw, editors. Biomimetic Neural Learning for Intelligent Robots: Intelligent Systems, Cognitive Robotics, and Neuroscience, volume 3575 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-27440-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TJ211.35 .I577 2004; Q334; TJ211.35 .I59 2004; Q334 .A48 2004; Internet; Q334 .A48 2004eb. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        27440-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3575.

Wirsing:2003:RTA

[WPH03] Martin Wirsing, Dirk Pattinson, and Rolf Hennicker, editors. Recent Trends in Algebraic Development Techniques: 16th International Workshop, WADT 2002, Frauenchiemsee, Germany, September 24–27, 2002: revised Selected Papers, volume 2755 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20537-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 W33 2002.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2755.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20537-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2755;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b94458.

Weyns:2005:EMA

[WPM05] Danny Weyns, H. Van Dyke Parunak, and Fabien Michel, editors. Environment for multi-agent systems: first international workshop, E4MAS, 2004, New York, NY, July 19, 2004. Revised selected papers, volume 3374 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-24575-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 E463 2004. URL http:

//www.springerlink.
com/content/978-3-540-
24575-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3374;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106134.

Weyns:2006:EMA

[WPM06] Danny Weyns, H. VanDyke Parunak, and Fabien Michel, editors. Environments for Multi-Agent Systems II: Second International Workshop, E4MAS 2005, Utrecht, The Netherlands, July 25, 2005, Selected Revised and Invited Papers, volume 3830 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32614-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32614-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3830.

Weyns:2007:EMA

[WPM07] Danny Weyns, H. Van Dyke Parunak, and Fabien Michel, editors. Environments for Multi-Agent Systems III: Third International Workshop, E4MAS 2006, Hakodate, Japan, May 8, 2006, Selected Revised and Invited Papers, volume 4389 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71102-3 (print), 3-540-71103-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71103-2.

Wang:2007:SPD

[WPR07] Qing Wang, Dietmar Pfahl, and David M. Raffo, editors. Software Process Dynamics and Agility: International Conference on Software Process, ICSP 2007, Minneapolis, MN, USA, May 19–20, 2007. Proceedings, volume 4470 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72425-7 (print), 3-540-72426-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72426-1.

Wang:2008:MGD

[WPR08] Qing Wang, Dietmar Pfahl, and David M. Raffo, editors. Making Globally Distributed Software Development a Success Story: International Conference on Software Process, ICSP 2008 Leipzig, Germany, May 10–11, 2008 Proceedings, volume 5007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-79587-1 (print), 3-540-79588-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        79588-9.

Wang:2006:SPC

[WPRW06] Qing Wang, Dietmar Pfahl, David M. Raffo, and Paul Wernick, editors. Software Process Change: International Software Process Workshop and International Workshop on Software Process Simulation and Modeling, SPW/ProSim 2006, Shanghai, China, May 20–21, 2006. Proceedings, volume 3966 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34199-4 (print), 3-540-34201-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        34201-4.

Wang:2006:RSK

[WPSY06] Guo-Ying Wang, James F. Peters, Andrzej Skowron, and Yiyu Yao, editors. Rough Sets and Knowledge Technology: First International Conference, RSKT 2006, Chongquing, China, July 24–26, 2006. Proceedings, volume 4062 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36297-5 (print), 3-540-36299-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36299-9.

Wang:2004:ISP

[WPV04] Huaxiong Wang, Josef Pieprzyk, and Vijay Varadharajan, editors. Information Security and Privacy: 9th Australasian Conference, ACISP 2004, Sydney, Australia, July 13–15, 2004: Proceedings, volume 3108 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22379-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I54 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3108.
htm;   http://www.springerlink.
com/content/978-3-540-
22379-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=

                                        0302-9743&volume=3108;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b98755.

Wolff:2011:KPD

[WPZA11] Karl Erich Wolff, Dmitry E. Palchunov, Nikolay G. Zagoruiko, and Urs Andelfinger, editors. Knowledge Processing and Data Analysis: First International Conference, KONT 2007, Novosibirsk, Russia, September 14–16, 2007 and First International Conference, KPP 2007,Darmstadt, Germany, September 28–30, 2007. Revised Selected Papers, volume 6581 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22139-4 (print), 3-642-22140-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        22140-8.

Wagner:2001:IAM

[WR01] Tom Wagner and Omer Rana, editors. Infrastructure for agents, multi-agent systems, and scalable multi-agent systems: International Workshop on Infrastructure for Scalable Multi-Agent Systems, Barcelona, Spain, June 3–7, 2000: Revised Papers, volume 1887 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-42315-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 I593 2000; QA267.A1 L43 no.1887. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1887.htm;
 http://www.springerlink.
com/content/978-3-540-
42315-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1887.

Windeatt:2003:MCS

[WR03] Terry Windeatt and Fabio Roli, editors. Multiple classifier systems: 4th International Workshop, MCS 2003, Guildford, UK, June 11– 13, 2003: Proceedings, volume 2709 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40369-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .M84 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2709.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        40369-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2709.

Weber:2007:CBR

[WR07] Rosina O. Weber and Michael M. Richter, editors. Case-Based Reasoning Research and Development: 7th International Conference on Case-Based Reasoning, ICCBR 2007 Belfast, Northern Ireland, UK, August 13–16, 2007 Proceedings, volume 4626 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74138-0 (print), 3-540-74141-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74141-1.

Wilkinson:2009:MMA

[WR09] Michael H. F. Wilkinson and Jos B. T. M. Roerdink, editors. Mathematical Morphology and Its Application to Signal and Image Processing: 9th International Symposium, ISMM 2009 Groningen, The Netherlands, August 24–27, 2009 Proceedings, volume 5720 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03612-0 (print), 3-642-03613-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
03613-2.

Wang:2011:AAA

[WR11] Dianhui Wang and Mark Reynolds, editors. AI 2011: Advances in Artificial Intelligence: 24th Australasian Joint Conference, Perth, Australia, December 5–8, 2011. Proceedings, volume 7106 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25831-X (print), 3-642-25832-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25832-9.

Wright:2003:FCI

[Wri03] Rebecca N. Wright, editor. Financial cryptography: 7th International Conference, FC 2003, Guadeloupe, French West Indies, January 27–30, 2003: revised papers, volume 2742 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40663-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HG1710 .F35 2003. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        40663-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2742;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b11831.

Wagner:2007:DES

[WRP07] Roland Wagner, Norman Revell, and Günther Pernul, editors. Database and Expert Systems Applications: 18th International Conference, DEXA 2007, Regensburg, Germany, September 3–7, 2007. Proceedings, volume 4653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74467-3 (print), 3-540-74469-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74469-6.

Weiss:1996:ALM

[WS96] Gerhard Weiss and Sandip Sen, editors. Adaption and learning in multiagent systems: IJCAI ’95 workshop, Montréal, Canada, August 21, 1995, proceedings, volume 1042 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60923-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .I55 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1042.
htm;   http://www.springerlink.
com/content/978-3-540-
60923-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1042.

Wermter:2000:HNS

[WS00] Stefan Wermter and Ron Sun. Hybrid neural systems, volume 1778 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67305-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 401 pp. LCCN QA76.87 .H79 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1778.
htm;   http://www.springerlink.
com/content/978-3-540-
67305-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1778.

Wachsmuth:2002:GSL

[WS02] Ipke Wachsmuth and Timo Sowa, editors. Gesture and sign language in human-computer interaction: international Gesture Workshop, GW 2001, London, UK, April 18–20, 2001: Revised Papers, volume 2298 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43678-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.H85 I618 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2298.htm;
http://www.springerlink.
com/content/978-3-540-
43678-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2298.

Williams:2006:DMT

[WS06] Graham J. Williams and Simeon J. Simoff, editors. Data Mining: Theory, Methodology, Techniques, and Applications, volume 3755 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32547-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
32547-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3755.

Wang:2012:ACA

[WS12] Xiaoyun Wang and Kazue Sako, editors. Advances in Cryptology — ASI-ACRYPT 2012: 18th International Conference on the Theory and Application of Cryptology and Information Security, Beijing, China, December 2–6, 2012. Proceedings, volume 7658 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34960-9 (print), 3-642-34961-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34961-4.

Wimmer:2008:EGI

[WSF08] Maria A. Wimmer, Hans J. Scholl, and Enrico Ferro, editors. Electronic Government: 7th International Conference, EGOV 2008, Turin, Italy, August 31 — September 5, 2008. Proceedings, volume 5184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85203-4 (print), 3-540-85204-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85204-9.

Wimmer:2007:EGI

[WSG07] Maria A. Wimmer, Jochen Scholl, and Åke Grönlund, editors. Electronic Government: 6th International Conference, EGOV 2007, Regensburg, Germany, September 3–7, 2007. Proceedings, volume 4656 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74443-6 (print), 3-540-74444-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74444-3.

Wimmer:2006:EGI

[WSGA06] Maria A. Wimmer, Hans J.

Scholl, Åke Grönlund, and Kim Viborg Andersen, editors. Electronic Government: 5th International Conference, EGOV 2006, Kraków, Poland, September 4– 8, 2006. Proceedings, volume 4084 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37686-0 (print), 3-540-37687-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37687-3.

Wimmer:2009:EGI

[WSJT09] Maria A. Wimmer, Hans J. Scholl, Marijn Janssen, and Roland Traunmüller, editors. Electronic Government: 8th International Conference, EGOV 2009, Linz, Austria, August 31 — September 3, 2009. Proceedings, volume 5693 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-03515-9 (print), 3-642-03516-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        03516-6.

Weber:1993:GDL

[WSL93] Matthias Weber, Martin Simons, and Christine Lafontaine. The generic development language Deva: presentation and case studies, volume 738 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 0-387-57335-6 (New York), 3-540-57335-6 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 246 pp. LCCN QA76.73.D4 W43 1993. DM58.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0738.htm;
http://www.springerlink.
com/content/978-0-387-
57335-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=738.

Washio:2006:NFA

[WSN^{+}06] Takashi Washio, Akito Sakurai, Katsuto Nakajima, Hideaki Takeda, Satoshi Tojo, and Makoto Yokoo, editors. New Frontiers in Artificial Intelligence: Joint JSAI 2005 Workshop Post-Proceedings, volume 4012 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35470-0 (print), 3-540-35471-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35471-0.

Wakamiya:2004:ANI

[WSS04] Naoki Wakamiya, Marcin Solarski, and James (P. G.) Sterbenz, editors. Active Networks: IFIP-TC6 5th International Working Conference, IWAN 2003 Kyoto, Japan, December 10–12, 2003 Proceedings, volume 2982 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-21250-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.5 .I95 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2982.htm;
 http://www.springerlink.
com/content/978-3-540-
21250-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2982;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b96396.

Washio:2007:NFA

[WSTI07] Takashi Washio, Ken Satoh, Hideaki Takeda, and Akihiro Inokuchi, editors. New Frontiers in Artificial Intelligence: JSAI 2006 Conference and Workshops, Tokyo, Japan, June 5– 9 2006, Revised Selected Papers, volume 4384 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-69901-5 (print), 3-540-69902-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69902-6.

Washio:2008:AKD

[WSTI08] Takashi Washio, Einoshin Suzuki, Kai Ming Ting, and Akihiro Inokuchi, editors. Advances in Knowledge Discovery and Data Mining: 12th Pacific-Asia Conference, PAKDD 2008 Osaka, Japan, May 20–23, 2008 Proceedings, volume 5012 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68124-8 (print), 3-540-68125-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68125-0.

Wagner:1996:DES

[WT96] Roland R. Wagner and Helmut Thoma, editors. Database and expert systems applications: 7th International Conference, DEXA ’96, Zürich, Switzerland, September 9–13, 1996: proceedings, volume 1134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61656-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3D29 1996.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1134.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        61656-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1134.

Ware:2007:WWG

[WT07] J. Mark Ware and George E. Taylor, editors. Web and Wireless Geographical Information Systems: 7th International Symposium, W2GIS 2007, Cardiff, UK, November 28–29, 2007. Proceedings, volume 4857 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76923-4 (print), 3-540-76925-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76925-5.

Wimmer:2005:EGI

[WTGA05] Maria A. Wimmer, Roland Traunmüller, Åke Grönlund, and Kim V. Andersen, editors. Electronic government: 4th international conference, EGOV 2005, Copenhagen, Denmark, August 22–26, 2005: proceedings, volume 3591 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-28466-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN JF1525.A8 E5843 2005. URL http://

springerlink.metapress.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3591;   http:
//www.springerlink.
com/content/978-3-540-
28466-6.

Widmayer:2002:ALP

[WTM^{+}02] Peter Widmayer, Francisco Triguero, Rafael Morales, Matthew Hennessy, Stephan Eidenbenz, and Ricardo Conejo, editors. Automata, Languages and Programming: 29th International Colloquium, ICALP 2002, Málaga, Spain, July 8–13, 2002. Proceedings, volume 2380 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43864-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I551 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2380.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43864-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2380.

Wiedermann:2006:STP

[WTP^{+}06] Jiří Wiedermann, Gerard Tel, Jaroslav Pokorný, Mária Bieliková, and Július Štuller, editors. SOFSEM 2006: Theory and Practice of Computer Science: 32nd Conference on Current Trends in Theory and Practice of Computer Science, Merin, Czech Republic, January 21– 27, 2006. Proceedings, volume 3831 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31198-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        31198-0;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3831.

Wang:2004:CMA

[WTZ^{+}04] Shan Wang, Katsumi Tanaka, Shuigeng Zhou, Tok Wang Ling, Jihong Guan, Dongqing Yang, Fabio Grandi, Eleni Mangina, Il-Yeol Song, and Heinrich C. Mayr, editors. Conceptual modeling for advanced application domains: ER 2004 workshops Comogis, Comwim, ECDM, Comoa, DGOV, and Ecomo, Shanghai, China, November 8–12, 2004, proceedings, volume 3289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23722-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
23722-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3289;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101694.

Wooldridge:1999:AIT

[WV99] Michael J. Wooldridge and Manuela M. Veloso, editors. Artificial Intelligence Today: Recent Trends and Developments, volume 1600 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66428-9. ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 487 pp. LCCN Q335 .A7871445 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1600.htm;
 http://www.springerlink.
com/content/978-3-540-
66428-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1600.

Wespi:2002:RAI

[WVD02] Andreas Wespi, Giovanni Vigna, and Luca Deri, editors. Recent Advances in Intrusion Detection: 5th International Symposium, RAID 2002, Zürich, Switzerland, October 16–18, 2002. Proceedings, volume 2516 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00020-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 R34 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2516.
htm;   http://www.springerlink.
com/content/978-3-540-
00020-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2516.

Westfechtel:2003:SCM

[WvdH03] Bernhard Westfechtel and Andre van der Hoek, editors. Software configuration management: ICSE Workshops SCM 2001 and SCM 2003, Toronto, Canada, May 14–15, 2001 and Portland, OR, USA, May 9– 10, 2003: Selected Papers, volume 2649 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-14036-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C69 I25 2001/03. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2649.htm;
http://www.springerlink.
com/content/978-3-540-
14036-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2649.

Wright:2005:FGA

[WVDS05] Alden H. Wright, Michael D.

Vose, Kenneth A. De Jong, and Lothar M. Schmitt, editors. Foundations of Genetic Algorithms: 8th International Workshop, FOGA 2005, Aizu-Wakamatsu City, Japan, January 5 — 9 , 2005, Revised Selected Papers, volume 3469 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27237-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA402.5 .W66 2005. URL http:

//www.springerlink.
com/content/978-3-540-
27237-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3469.

Wells:2002:EPA

[WW02] Don Wells and Laurie Williams, editors. Extreme Programming and Agile Methods— XP/Agile Universe 2002: Second XP Universe and First Agile Universe Conference, Chicago, IL, USA, August 4–7, 2002. Proceedings, volume 2418 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44024-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 X63 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2418.htm;
 http://www.springerlink.
com/content/978-3-540-
44024-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2418.

Watson:2003:IAA

[WW03] Bruce William Watson and Derick Wood, editors. Implementation and application of automata: 6th International Conference, CIAA 2001, Pretoria, South Africa, July 23–25, 2001: Revised Papers, volume 2494 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-00400-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .I5787 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2494.htm;
 http://www.springerlink.
com/content/978-3-540-
00400-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2494.

Also available via the World Wide Web.

Wagner:2007:ASA

[WW07] Dorothea Wagner and Roger Wattenhofer, editors. Algorithms for Sensor and Ad Hoc Networks: Advanced Lectures, volume 4621 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74990-X (print), 3-540-74991-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        74991-2.

Wooldridge:2002:AOS

[WWC02] Michael J. Wooldridge, Gerhard Weiss, and Paolo Ciancarini, editors. Agent-oriented software engineering II: Second International Workshop, AOSE 2001, Montréal, Canada, May 29, 2001: revised papers and invited contributions, volume 2222 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-43282-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .A57 2001. URL http:

                                        //link.springer-ny.

com/link/service/series/
0558/tocs/t2222.htm;
http://www.springerlink.
com/content/978-3-540-
43282-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2222.

Wing:1999:FFMa

[WWD99a] Jeannette Marie Wing, Jim C. P. Woodcock, and Jim Davies, editors. FM’99 — formal methods: World Congress on Formal Methods in the Development of Computing Systems, Toulouse, France, September 20– 24, 1999: Proceedings, volume 1708 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66587-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1708-1709. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1708.
htm;   http://www.springerlink.
com/content/978-3-540-
66587-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1708.

Wing:1999:FFMb

[WWD99b] Jeannette Marie Wing, Jim C. P. Woodcock, and Jim Davies, editors. FM’99 — formal methods: World Congress on Formal Methods in the Development of Computing Systems, Toulouse, France, September 20– 24, 1999: Proceedings, volume 1709 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66588-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1708-1709. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t1709.
htm;   http://www.springerlink.
com/content/978-3-540-
66588-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1709.

Wood:1998:AIS

[WY98] Derick Wood and Sheng Yu, editors. Automata implementation: Second International Workshop on Implementing Automata, WIA ’97, London, Ontario, Canada, September 18–20, 1997: Revised Papers, volume 1436 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64694-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1436. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1436.htm;
 http://www.springerlink.
com/content/978-3-540-
64694-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1436.

Watanabe:1999:ALT

[WY99] Osamu Watanabe and Takashi Yokomori, editors. Algorithmic learning theory: 10th International Conference, ALT’99, Tokyo, Japan, December 6–8, 1999: Proceedings, volume 1720 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66748-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1720.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1720.
htm;   http://www.springerlink.
com/content/978-3-540-
66748-3;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=1720.

Webb:2004:AAA

[WY04] Geoffrey I. Webb and Xinghuo Yu, editors. AI 2004: Advances in Artificial Intelligence: 17th Australian Joint Conference on Artificial Intelligence, Cairns, Australia, December 4–6, 2004: Proceedings, volume 3339 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24059-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .L48 no. 3339. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        24059-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3339;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104336.

Wang:2001:AWA

[WYL01] Xiaoyang Sean Wang, Ge Yu, and Hongjun Lu, editors. Advances in Web-age information management: First International Conference, WAIM 2001, Xi’an, China, July 9–11, 2001: Proceedings, volume 2118 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42298-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 W325 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2118.htm;
 http://www.springerlink.
com/content/978-3-540-
42298-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2118.

Wang:2012:ANNa

[WYP12a] Jun Wang, Gary G. Yen, and Marios M. Polycarpou, editors. Advances in Neural Networks — ISNN 2012: 9th International Symposium on Neural Networks, Shenyang, China, July 11–14, 2012. Proceedings, Part I, volume 7367 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31345-0 (print), 3-642-31346-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        31346-2.

Wang:2012:ANNb

[WYP12b] Jun Wang, Gary G. Yen, and Marios M. Polycarpou, editors. Advances in Neural Networks — ISNN 2012: 9th International Symposium on Neural Networks, Shenyang, China, July 11–14, 2012. Proceedings, Part II, volume 7368 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31361-2 (print), 3-642-31362-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31362-2.

Wang:2010:MLM

[WYSS10] Fei Wang, Pingkun Yan, Kenji Suzuki, and Dinggang Shen, editors. Machine Learning in Medical Imaging: First International Workshop, MLMI 2010, Held in Conjunction with MICCAI 2010, Beijing, China, September 20, 2010. Proceedings, volume 6357 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15947-8 (print), 3-642-15948-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15948-0.

Wu:2008:AGP

[WYX08] Song Wu, Laurence T.

Yang, and Tony Li Xu, editors. Advances in Grid and Pervasive Computing: Third International Conference, GPC 2008, Kunming, China, May 25–28, 2008. Proceedings, volume 5036 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68081-0 (print), 3-540-68083-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
68083-3.

Wang:2006:ANNa

[WYZ06a] Jun Wang, Zhang Yi, and Jacek M. Zurada, editors. Advances in Neural Networks: ISNN 2006: Third International Symposium on Neural Networks, ISNN 2006, Chengdu, China, May 28: June 1, 2006, Proceedings, Part I, volume 3971–3973 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34439-X (part 1), 3-540-34437-3 (part 2), 3-540-34482-9 (part 3). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34439-1.

Wang:2006:ANNb

[WYZ^{+}06b] Jun Wang, Zhang Yi, Jacek M. Zurada, Bao-Liang Lu, and Hujun Yin, editors. Advances in Neural Networks — ISNN 2006: Third International Symposium on Neural Networks, Chengdu, China, May 28 — June 1, 2006, Proceedings, Part II, volume 3972 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34437-3 (print), 3-540-34438-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34438-4.

Warnow:2003:CCA

[WZ03] Tandy Warnow and Binhai Zhu, editors. Computing and combinatorics: 9th annual International Conference, COCOON 2003, Big Sky, MT, USA, July 25–28, 2003: Proceedings, volume 2697 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40534-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .C549 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2697.htm;
http://www.springerlink.
com/content/978-3-540-
40534-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2697.

Wobcke:2008:AAA

[WZ08] Wayne Wobcke and Mengjie Zhang, editors. AI 2008: Advances in Artificial Intelligence: 21st Australasian Joint Conference on Artificial Intelligence Auckland, New Zealand, December 1–5, 2008. Proceedings, volume 5360 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89377-6 (print), 3-540-89378-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
89378-3.

Watanabe:2009:SAF

[WZ09] Osamu Watanabe and Thomas Zeugmann, editors. Stochastic Algorithms: Foundations and Applications: 5th International Symposium, SAGA 2009, Sapporo, Japan, October 26–28, 2009. Proceedings, volume 5792 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04943-5 (print), 3-642-04944-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04944-6.

Wolff:2011:TSS

[WZ11] Burkhart Wolff and Fatiha Zaïdi, editors. Testing Software and Systems: 23rd IFIP WG 6.1 International Conference, ICTSS 2011, Paris, France, November 7–10, 2011. Proceedings, volume 7019 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24579-X (print), 3-642-24580-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24580-0.

Winter:2010:GPD

[WZA10] Robert Winter, J. Leon Zhao, and Stephan Aier, editors. Global Perspectives on Design Science Research: 5th International Conference, DESRIST 2010, St. Gallen, Switzerland, June 4–5, 2010. Proceedings., volume 6105 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13334-7 (print), 3-642-13335-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13335-0.

Wang:2011:COA

[WZD11] Weifan Wang, Xuding Zhu, and Ding-Zhu Du, editors. Combinatorial Optimization and Applications: 5th International Conference, CO-COA 2011, Zhangjiajie, China, August 4–6, 2011. Proceedings, volume 6831 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-22615-9 (print), 3-642-22616-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
22616-8.

Washio:2007:ETK

[WZH^{+}07] Takashi Washio, Zhi-Hua Zhou, Joshua Zhexue Huang, Xiaohua Hu, Jinyan Li, Chao Xie, Jieyue He, Deqing Zou, Kuan-Ching Li, and Mário M. Freire, editors. Emerging Technologies in Knowledge Discovery and Data Mining: PAKDD 2007 International Workshops Nanjing, China, May 22– 25, 2007 Revised Selected Papers, volume 4819 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77016-X (print), 3-540-77018-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77018-3.

Wang:2012:WTA

[WZH^{+}12] Hua Wang, Lei Zou, Guangyan Huang, Jing He, Chaoyi Pang, Hao Lan Zhang, Dongyan Zhao, and Zhuang Yi, editors. Web Technologies and Applications: AP-Web 2012 International Workshops: SenDe, IDP, IEKB, MBC, Kunming, China, April 11– 13, 2012. Proceedings, volume 7234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29425-1 (print), 3-642-29426-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29426-6.

Wang:2012:WAS

[WZJX12] Xinbing Wang, Rong Zheng, Tao Jing, and Kai Xing, editors. Wireless Algorithms, Systems, and Applications: 7th International Conference, WASA 2012, Yellow Mountains, China, August 8–10, 2012. Proceedings, volume 7405 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31868-1 (print), 3-642-31869-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31869-6.

Xiang:2012:DKE

[X^{+}12a] Yang Xiang et al., editors. Data and knowledge engineering: third International Conference, ICDKE 2012, Wuyishan, Fujian, China, November 21–23, 2012. Proceedings, volume 7696 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34678-2, 3-642-34679-0 (e-book). LCCN QA76.9.D3 I58 2012.

                                        URL http://www.springerlink.
                                        com/content/978-3-642-
                                        34678-1;   http://www.
                                        springerlink.com/content/
                                        978-3-642-34679-8.

Xiang:2012:IDC

[X^{+}12b] Yang Xiang et al., editors. Internet and distributed computing systems: 5th International Conference, IDCS 2012, Wuyishan, Fujian, China, November 21–23, 2012. Proceedings, volume 7646 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 3-642-34882-3, 3-642-34883-1 (e-book). LCCN TK5105.875.I57 .I58 2012.

URL http://www.springerlink.
com/content/978-3-642-
34882-2;   http://www.
springerlink.com/content/
978-3-642-34883-9.

Xu:2012:NSS

[XBM12] Li Xu, Elisa Bertino, and Yi Mu, editors. Network and System Security: 6th International Conference, NSS 2012, Wuyishan, Fujian, China, November 21–23, 2012. Proceedings, volume 7645 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34600-6 (print), 3-642-34601-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
34601-9.

Xie:2010:ATC

[XBS^{+}10] Bing Xie, Juergen Branke, S. Masoud Sadjadi, Daqing Zhang, and Xingshe Zhou, editors. Autonomic and Trusted Computing: 7th International Conference, ATC 2010, Xi’an, China, October 26–29, 2010. Proceedings, volume 6407 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16575-3 (print), 3-642-16576-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16576-4.

Xiang:2003:AAI

[XCd03] Yang Xiang and Brahim Chaib-draa, editors. Advances in artificial intelligence: 16th Conference of the Canadian Society for Computational Studies of Intelligence, AI 2003, Halifax, Canada, June 2003: Proceedings, volume 2671 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40300-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .C36 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2671.htm;
http://www.springerlink.
com/content/978-3-540-
40300-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2671.

Xiang:2011:AAPb

[XCHZ11a] Yang Xiang, Alfredo Cuzzocrea, Michael Hobbs, and Wanlei Zhou, editors. Algorithms and Architectures for Parallel Processing: 11th International Conference, ICA300 2011, Melbourne, Australia, October 24–26, 2011, Proceedings, Part II, volume 7017 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24668-0 (print), 3-642-24669-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24669-2.

Xiang:2011:AAPa

[XCHZ11b] Yang Xiang, Alfredo Cuzzocrea, Michael Hobbs, and Wanlei Zhou, editors. Algorithms and Architectures for Parallel Processing: 11th International Conference, ICA3PP, Melbourne, Australia, October 24–26, 2011, Proceedings, Part I, volume 7016 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24649-4 (print), 3-642-24650-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24650-0.

Xing:2011:DLC

[XCR11] Chunxiao Xing, Fabio Crestani, and Andreas Rauber, editors. Digital Libraries: For Cultural Heritage, Knowledge Dissemination, and Future Creation: 13th International Conference on Asia-Pacific Digital Libraries, ICADL 2011, Beijing, China, October 24–27, 2011. Proceedings, volume 7008 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24825-X (print), 3-642-24826-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
24826-9.

Xiong:2008:IRAa

[XHXL08] Caihua Xiong, Yongan Huang, Youlun Xiong, and Honghai Liu, editors. Intelligent Robotics and Applications: First International Conference, ICIRA 2008, Wuhan, China, October 15–17, 2008, Proceedings, Part I, volume 5314 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88512-9 (print), 3-540-88513-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88513-9.

Xiao:2012:GIS

[XKGS12] Ningchuan Xiao, Mei-Po Kwan, Michael F. Goodchild, and Shashi Shekhar, editors. Geographic Information Science: 7th International Conference, GIScience 2012, Columbus, OH, USA, September 18– 21, 2012. Proceedings, volume 7478 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33023-1 (print), 3-642-33024-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33024-7.

Xiong:2011:KSE

[XL11] Hui Xiong and W. B. Lee, editors. Knowledge Science, Engineering and Management: 5th International Conference, KSEM 2011, Irvine, CA, USA, December 12–14, 2011. Proceedings, volume 7091 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25974-X (print), 3-642-25975-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25975-3.

Xiong:2008:IRAb

[XLHX08] Caihua Xiong, Honghai Liu, Yongan Huang, and Youlun Xiong, editors. Intelligent Robotics and Applications: First International Conference, ICIRA 2008 Wuhan, China, October 15–17, 2008 Proceedings, Part II, volume 5315 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88516-1 (print), 3-540-88518-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88518-4.

Xiang:2012:CSS

[XLKZ12] Yang Xiang, Javier Lopez, C.-C. Jay Kuo, and Wanlei Zhou, editors. Cyberspace Safety and Security: 4th International Symposium, CSS 2012, Melbourne, Australia, December 12– 13, 2012. Proceedings, volume 7672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35361-4 (print), 3-642-35362-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35362-8.

Xiang:2012:AAPa

[XSA^{+}12a] Yang Xiang, Ivan Stojmenovic, Bernady O. Apduhan, Guojun Wang, Koji Nakano, and Albert Zomaya, editors. Algorithms and Architectures for Parallel Processing: 12th International Conference, ICA3PP 2012, Fukuoka, Japan, September 4–7, 2012, Proceedings, Part I, volume 7439 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33077-0 (print), 3-642-33078-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33078-0.

Xiang:2012:AAPb

[XSA^{+}12b] Yang Xiang, Ivan Stojmenovic, Bernady O. Apduhan, Guojun Wang, Koji Nakano, and Albert Zomaya, editors. Algorithms and Architectures for Parallel Processing: 12th International Conference, ICA3PP 2012, Fukuoka, Japan, September 4–7, 2012, Proceedings, Part II, volume 7440 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33064-9 (print), 3-642-33065-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
33065-0.

Xenakis:2011:CII

[XW11] Christos Xenakis and Stephen Wolthusen, editors. Critical Information Infrastructures Security: 5th International Workshop, CRITIS 2010, Athens, Greece, September 23– 24, 2010. Revised Papers, volume 6712 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21693-5 (print), 3-642-21694-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
21694-7.

Xie:2009:IRA

[XXX^{+}09] Ming Xie, Youlun Xiong, Caihua Xiong, Honghai Liu, and Zhencheng Hu, editors. Intelligent Robotics and Applications: Second International Conference, ICIRA 2009, Singapore, December 16–18, 2009. Proceedings, volume 5928 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10816-4 (print), 3-642-10817-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        10817-4.

Xiao:2007:ATC

[XYM^{+}07] Bin Xiao, Laurence T. Yang, Jianhua Ma, Christian Muller-Schloer, and Yu Hua, editors. Autonomic and Trusted Computing: 4th International Conference, ATC 2007, Hong Kong, China, July 11–13, 2007. Proceedings, volume 4610 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-73546-1 (print), 3-540-73547-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        73547-2.

Xu:2011:DSA

[XYZU11] Jianliang Xu, Ge Yu, Shuigeng Zhou, and Rainer Unland, editors. Database Systems for Adanced Applications: 16th International Conference, DASFAA 2011, International Workshops: GDB, SIM3, FlashDB, SNSMW, DaMEN, DQIS, Hong Kong, China, April 22–25, 2011. Proceedings, volume 6637 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20243-8 (print), 3-642-20244-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20244-5.

Xu:2007:APP

[XZCL07] Ming Xu, Yinwei Zhan, Jiannong Cao, and Yijun Liu, editors. Advanced Parallel Processing Technologies: 7th International Symposium, APPT 2007 Guangzhou, China, November 22– 23, 2007 Proceedings, volume 4847 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76836-X (print), 3-540-76837-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76837-1.

Yin:2002:IDE

[YAF^{+}02] Hujun Yin, Nigel Allinson, Richard Freeman, John Keane, and Simon Hubbard, editors. Intelligent Data Engineering and Automated Learning— IDEAL 2002: Third International Conference, Manchester, UK, August 12–14,2002. Proceedings, volume 2412 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-44025-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55864 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2412.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        44025-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2412.

Yahav:2011:SAI

[Yah11] Eran Yahav, editor. Static Analysis: 18th International Symposium, SAS 2011, Venice, Italy, September 14–16, 2011. Proceedings, volume 6887 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23701-0 (print), 3-642-23702-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23702-7.

Yakhno:2000:AIS

[Yak00] Tatyana Yakhno, editor.

Advances in Information Systems: First International Conference, AD-VIS 2000, Izmir, Turkey, October 25–27, 2000. Proceedings, volume 1909 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41184-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T58.5 .I566 2000.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1909.
htm;   http://www.springerlink.
com/content/978-3-540-
41184-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1909.

Yakhno:2002:AIS

[Yak02] Tatyana Yakhno, editor. Advances in Information Systems: Second International Conference, ADVIS 2002, Izmir, Turkey, October 23–25, 2002. Proceedings, volume 2457 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-00009-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T58.5 .I566 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2457.htm;
 http://www.springerlink.
com/content/978-3-540-
00009-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2457.

Yakhno:2004:AIS

[Yak04] Tatyana Yakhno, editor. Advances in Information Systems: Third International Conference, AD-VIS 2004, Izmir, Turkey, October 20–22, 2004, Proceedings, volume 3261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23478-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3261.htm;

http://www.springerlink.
com/content/978-3-540-
23478-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3261;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b101594.

Ystad:2012:SSM

[YAKM^{+}12] Sølvi Ystad, Mitsuko Aramaki, Richard Kronland-Martinet, Kristoffer Jensen, and Sanghamitra Mohanty, editors. Speech, Sound and Music Processing: Embracing Research in India: 8th International Symposium, CMMR 2011, 20th International Symposium, FRSM 2011, Bhubaneswar, India, March 9–12, 2011, Revised Selected Papers, volume 7172 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31979-3 (print), 3-642-31980-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
31980-8.

Ystad:2010:ADI

[YAKMJ10] Sølvi Ystad, Mitsuko Aramaki, Richard Kronland-Martinet, and Kristoffer Jensen, editors. Auditory Display: 6th International Symposium, CMMR/ICAD 2009, Copenhagen, Denmark, May 18–22, 2009. Revised Papers, volume 5954 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12438-0 (print), 3-642-12439-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12439-6.

Ystad:2011:EMC

[YAKMJ11] Sølvi Ystad, Mitsuko Aramaki, Richard Kronland-Martinet, and Kristoffer Jensen, editors. Exploring Music Contents: 7th International Symposium, CMMR 2010, Málaga, Spain, June 21– 24, 2010. Revised Papers, volume 6684 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23125-X (print), 3-642-23126-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23126-1.

Yang:2005:EUC

[YAL^{+}05] Laurence Tianruo Yang, Makoto Amamiya, Zhen Liu, Minyi Guo, and Franz J. Rammig, editors. Embedded and ubiquitous computing–EUC 2005: international conference, EUC 2005, Nagasaki, Japan, December 6–9, 2005: proceedings, volume 3824 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30807-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 E962 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30807-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3824.

Yalcin:2010:RFI

[Yal10] Siddika Berna Ors Yalcin, editor. Radio Frequency Identification: Security and Privacy Issues: 6th International Workshop, RFID-Sec 2010, Istanbul, Turkey, June 8–9, 2010, Revised Selected Papers, volume 6370 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16821-3 (print), 3-642-16822-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16822-2.

Yamaguchi:2008:PAK

[Yam08] Takahira Yamaguchi, editor. Practical Aspects of Knowledge Management: 7th International Conference, PAKM 2008, Yokohama, Japan, November 22–23, 2008. Proceedings, volume 5345 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89446-2 (print), 3-540-89447-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89447-6.

Yang:2011:PQC

[Yan11a] Bo-Yin Yang, editor. Post-Quantum Cryptography: 4th International Workshop, PQCrypto 2011, Taipei, Taiwan, November 29 — December 2, 2011. Proceedings, volume 7071 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25404-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25404-8.

Yang:2011:PLS

[Yan11b] Hongseok Yang, editor. Programming Languages and Systems: 9th Asian Symposium, APLAS 2011, Kenting, Taiwan, December 5– 7, 2011. Proceedings, volume 7078 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25317-2 (print), 3-642-25318-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25318-8.

Yao:1982:DBDa

[Yao82] S. Bing Yao, editor. Data base design techniques I: requirements and logical structures, NYU symposium, New York, May 1978, volume 132 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CO-DEN LNCSD9. ISBN 0-387-11214-6 (paperback: U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA 76.9 D3 D293 1978. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0132.htm;
 http://www.springerlink.
com/content/978-0-387-
11214-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=132.

Yao:1995:PEC

[Yao95] Xin Yao, editor. Progress in evolutionary computation: AI’93 and AI’94 Workshops on Evolutionary Computation, Melbourne, Victoria, Australia, November 16, 1993, Armidale, NSW, Australia, November 21– 22, 1994: Selected Papers, volume 956 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60154-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.618 .A2 1993. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0956.htm;
http://www.springerlink.
com/content/978-3-540-
60154-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=956.

Yasuda:2000:ANS

[Yas00] Hiroshi Yasuda, editor.

Active Networks: Second International Working Conference, IWAN 2000, Tokyo, Japan, October 16–18, 2000. Proceedings, volume 1942 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41179-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1942; TK5105.5 .I95 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1942.htm;
http://www.springerlink.
com/content/978-3-540-
41179-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1942.

Yao:2004:PPS

[YBL^{+}04] Xin Yao, Edmund Burke, José A. Lozano, Jim Smith, Juan J. Merelo-Guervós, John A. Bullinaria, Jonathan Rowe, Peter Tiño, Ata Kabán, and Hans-Paul Schwefel, editors. Parallel problem solving from nature— PPSN VIII: 8th international conference, Birmingham, UK, September 18–22, 2004: proceedings, volume 3242 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23092-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .C69 2004. URL http:

//www.springerlink.
com/content/978-3-540-
23092-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3242;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100601.

Yager:1987:UKB

[YBM87] Ronald R. Yager and B. Bouchon-Meunier, editors. Uncertainty in knowledge-based systems: International Conference on Information Processing and Management of Uncertainty in Knowledge-Based Systems, Paris, France, June 30–July 4, 1986:selected and extended contributions, volume 286 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 3-540-18579-8 (Berlin), 0-387-18579-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I5741 1986. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0286.htm;
 http://www.springerlink.
com/content/978-3-540-
18579-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=286.

Yoshida:2011:VCA

[YC11] Hiroyuki Yoshida and Wenli Cai, editors. Virtual Colonoscopy and Abdominal Imaging. Computational Challenges and Clinical Opportunities: Second International Workshop, Held in Conjunction with MICCAI 2010, Beijing, China, September 20, 2010, Revised Selected Papers, volume 6668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25718-6 (print), 3-642-25719-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25719-3.

Yin:2012:IDE

[YCB12] Hujun Yin, José A. F. Costa, and Guilherme Barreto, editors. Intelligent Data Engineering and Automated Learning — IDEAL 2012: 13th International Conference, Natal, Brazil, August 29–31, 2012. Proceedings, volume 7435 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32638-2 (print), 3-642-32639-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        32639-4.

Yang:2008:ISI

[YCC^{+}08] Christopher C. Yang, Hsinchun Chen, Michael Chau, Kuiyu Chang, Sheau-Dong Lang, Patrick S. Chen, Raymond Hsieh, Daniel Zeng, Fei-Yue Wang, Kathleen Carley, Wenji Mao, and Justin Zhan, editors. Intelligence and Security Informatics: IEEE ISI 2008 International Workshops: PAISI, PACCF, and SOCO 2008, Taipei, Taiwan, June 17, 2008. Proceedings, volume 5075 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-69136-7 (print), 3-540-69304-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
69304-8.

Cai:2007:TAM

[yCCZ07] Jin yi Cai, S. B.

(S. Barry) Cooper, and Hong Zhu, editors. Theory and applications of models of computation: 4th international conference, TAMC 2007, Shanghai, China, May 22–25, 2007: proceedings, volume 4484 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. ISBN 3-540-72503-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.7 .T36 2007. URL http:

//www.springerlink.
com/content/978-3-540-
72503-9;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=4484.

Cai:1996:CCS

[yCW96] Jin yi Cai and Chak Kuen Wong, editors. Computing and combinatorics: second annual international conference, CO-COON ’96, Hong Kong, June 17–19, 1996: proceedings, volume 1090 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61332-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I586 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1090.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        61332-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1090.

Yalamanchili:1998:PCR

[YD98] Sudhakar Yalamanchili and Jose Duato, editors. Parallel computer routing and communication: Second International Workshop, PCRCW 97, Atlanta, Georgia, USA, June 26–27, 1997: Proceedings, volume 1417 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-64571-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1417.

URL http://lcweb.
loc.gov/catdir/toc/
98-24763.html;   http:
//link.springer-ny.
com/link/service/series/
0558/tocs/t1417.htm;
 http://www.springerlink.
com/content/978-3-540-
64571-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1417.

Yung:2006:PKC

[YDKM06] Moti Yung, Yevgeniy Dodis, Aggelos Kiayias, and Tal Malkin, editors. Public Key Cryptography: PPK 2006: 9th International Conference on Theory and Practice in Public-Key Cryptography, New York, NY, USA, April 24–26, 2006. Proceedings, volume 3958 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33851-9 (softcover). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33851-2;   http://www.
springerlink.com/openurl.

                                        asp?genre=issue&issn=
                                        0302-9743&volume=3958.

Yang:2008:COA

[YDW08] Boting Yang, Ding-Zhu Du, and Cao An Wang, editors. Combinatorial Optimization and Applications: Second International Conference, CO-COA 2008, St. John’s, NL, Canada, August 21– 24, 2008. Proceedings, volume 5165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85096-1 (print), 3-540-85097-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        85097-7.

Yellin:1988:AGI

[Yel88] Daniel M. Yellin. Attribute grammar inversion and source-to-source translation, volume 302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1988. CO-DEN LNCSD9. ISBN 3-540-19072-4 (Berlin), 0-387-19072-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 176 pp. LCCN QA76.76.T83 Y451 1988.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0302.
htm;   http://www.springerlink.
com/content/978-3-540-
19072-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=302.

Yang:2004:IDE

[YEY04] Zhen Rong Yang, Richard Everson, and Hujun Yin, editors. Intelligent Data Engineering and Automated Learning—IDEAL 2004: 5th International Conference, Exeter, UK, August 25–27, 2004: Proceedings, volume 3177 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22881-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3177.htm;
 http://www.springerlink.
com/content/978-3-540-
22881-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3177;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99975.

Feng:1975:PPP

[yF75] Tse yun Feng, editor. Parallel processing: proceedings of the Sagamore Computer Conference, August 20–23, 1974, volume 24 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1975. CODEN LNCSD9. ISBN 0-387-07135-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 S231 1974. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0024.htm;
                                         http://www.springerlink.
                                        com/content/978-0-387-
                                        07135-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=24.

Yang:2012:SCB

[YGE12] Shanchieh Jay Yang, Ariel M. Greenberg, and Mica Endsley, editors. Social Computing, Behavioral — Cultural Modeling and Prediction: 5th International Conference, SBP 2012, College Park, MD, USA, April 3–5, 2012. Proceedings, volume 7227 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29046-9 (print), 3-642-29047-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
29047-3.

Yang:2004:EUC

[YGGJ04] Laurence T. Yang, Minyi Guo, Guang R. Gao, and Niraj K. Jha, editors. Embedded and Ubiquitous Computing: International Conference EUC 2004, Aizu-Wakamatsu City, Japan, August 25–27, 2004, volume 3207 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22906-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3207.htm;
http://www.springerlink.
com/content/978-3-540-
22906-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3207;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100039.

Yolum:2005:CIS

[YGGÖ05] Pinar Yolum, Tunga Güngör, Fikret Gürgen, and Can Özturan, editors. Computer and Information Sciences — ISCIS 2005: 20th International Symposium, Istanbul, Turkey, October 26–28, 2005. Proceedings, volume 3733 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29414-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29414-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3733.

Yu:2010:RSK

[YGL^{+}10] Jian Yu, Salvatore Greco, Pawan Lingras, Guoyin Wang, and Andrzej Skowron, editors. Rough Set and Knowledge Technology: 5th International Conference, RSKT 2010, Beijing, China, October 15–17, 2010. Proceedings, volume 6401 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16247-9 (print), 3-642-16248-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
16248-0.

Yang:2009:MICa

[YHR^{+}09a] Guang-Zhong Yang, David Hawkes, Daniel Rueckert, Alison Noble, and Chris Taylor, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2009: 12th International Conference, London, UK, September 20–24, 2009, Proceedings, Part I, volume 5761 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04267-8 (print), 3-642-04268-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04268-3.

Yang:2009:MICb

[YHR^{+}09b] Guang-Zhong Yang, David Hawkes, Daniel Rueckert, Alison Noble, and Chris Taylor, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2009: 12th International Conference, London, UK, September 20–24, 2009, Proceedings, Part II, volume 5762 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04270-8 (print), 3-642-04271-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04271-3.

Yoshida:2012:AICb

[YHV12] Hiroyuki Yoshida, David Hawkes, and Michael W. Vannier, editors. Abdominal Imaging. Computational and Clinical Applications: 4th International Workshop, Held in Conjunction with MIC-CAI 2012, Nice, France, October 1, 2012. Proceedings, volume 7601 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33611-6 (print), 3-642-33612-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-642-
33612-6.

Yu:2009:ANNa

[YHZ09a] Wen Yu, Haibo He, and Nian Zhang, editors. Advances in Neural Networks — ISNN 2009: 6th International Symposium on Neural Networks, ISNN 2009 Wuhan, China, May 26– 29, 2009 Proceedings, Part I, volume 5551 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01506-9 (print), 3-642-01507-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01507-6.

Yu:2009:ANNb

[YHZ09b] Wen Yu, Haibo He, and Nian Zhang, editors. Advances in Neural Networks — ISNN 2009: 6th International Symposium on Neural Networks, ISNN 2009 Wuhan, China, May 26– 29, 2009 Proceedings, Part II, volume 5552 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01509-3 (print), 3-642-01510-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01510-6.

Yu:2009:ANNc

[YHZ09c] Wen Yu, Haibo He, and Nian Zhang, editors. Advances in Neural Networks — ISNN 2009: 6th International Symposium on Neural Networks, ISNN 2009 Wuhan, China, May 26– 29, 2009 Proceedings, Part III, volume 5553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01512-3 (print), 3-642-01513-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        01513-7.

Yonezawa:1991:CTL

[YI91] Akinori Yonezawa and T. Ito, editors. Concurrency: theory, language, and architecture: UK/ Japan workshop, Oxford, UK, September 25–27, 1989: proceedings, volume 491 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CO-DEN LNCSD9. ISBN 3-540-53932-8 (Germany), 0-387-53932-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .U5 1989. DM49.50. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t0491.
htm;   http://www.springerlink.
com/content/978-3-540-
53932-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=491.

Yi:2005:PLS

[Yi05] Kwangkeun Yi, editor. Programming languages and systems: third Asian symposium, APLAS 2005, Tsukuba, Japan, November 2– 5, 2005: proceedings, volume 3780 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29735-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .A65 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29735-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3780.

Yi:2006:SAI

[Yi06] Kwangkeun Yi, editor. Static Analysis: 13th International Symposium, SAS 2006, Seoul, Korea, August 29–31, 2006. Proceedings, volume 4134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37756-5 (print), 3-540-37758-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37758-0.

Yen:2012:DLT

[YI12] Hsu-Chun Yen and Oscar H. Ibarra, editors. Developments in Language Theory: 16th International Conference, DLT 2012, Taipei, Taiwan, August 14–17, 2012. Proceedings, volume 7410 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31652-2 (print), 3-642-31653-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.

com/content/978-3-642-
31653-1.

Yang:2004:MIA

[YJ04] Guang-Zhong Yang and Tianzi Jiang, editors. Medical Imaging and Augmented Reality: Second International Workshop, MIAR 2004, Beijing, China, August 19– 20, 2004. Proceedings, volume 3150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22877-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3150.htm;
http://www.springerlink.
com/content/978-3-540-
22877-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3150;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99698.

Yli-Jyra:2006:FSM

[YJKK06] Anssi Yli-Jyrä, Lauri Karttunen, and Juhani Karhumäki, editors. Finite-State Methods and Natural Language Processing: 5th International Workshop, FSMNLP 2005, Helsinki, Finland, September 1–2, 2005. Revised Papers, volume 4002 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35467-0 (print), 3-540-35469-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35469-7.

Yli-Jyra:2010:FSM

[YJKSW10] Anssi Yli-Jyrä, András Kornai, Jacques Sakarovitch, and Bruce Watson, editors. Finite-State Methods and Natural Language Processing: 8th International Workshop, FSMNLP 2009, Pretoria, South Africa, July 21–24, 2009, Revised Selected Papers, volume 6062 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14683-X (print), 3-642-14684-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14684-8.

Yang:2006:ATC

[YJMU06] Laurence T. Yang, Hai Jin, Jianhua Ma, and Theo Ungerer, editors. Autonomic and Trusted Computing: Third International Conference, ATC 2006, Wuhan, China, September 3– 6, 2006. Proceedings, volume 4158 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-38619-X (print), 3-540-38622-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
38622-3.

Yang:2006:MIA

[YJS^{+}06] Guang-Zhong Yang, TianZi Jiang, Dinggang Shen, Lixu Gu, and Jie Yang, editors. Medical Imaging and Augmented Reality: Third International Workshop, Shanghai, China, August 17– 18, 2006 Proceedings, volume 4091 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37220-2 (print), 3-540-37221-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37221-9.

Yao:1982:DBDb

[YK82] S. Bing Yao and Toshiyasu Kunii, editors. Data base design techniques II: physical structures and applications proceedings, Tokyo, November 1979, volume 133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1982. CODEN LNCSD9. ISBN 3-540-11215-4, 0-387-11215-

4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9 .D3D322. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0133.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        11215-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=133.

Yao:1997:SEL

[YKF97] Xin Yao, Jong-Hwan Kim, and T. (Takeshi) Furuhashi, editors. Simulated evolution and learning: First Asia-Pacific Conference, SEAL ’96, Taejon, Korea, November 1–9, 1996:

Selected Papers, volume 1285 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63399-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C65S42 1996. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1285.htm;
 http://www.springerlink.
com/content/978-3-540-
63399-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1285.

Yeung:2006:SSS

[YKF^{+}06] Dit-Yan Yeung, James T.

Kwok, Ana Fred, Fabio Roli, and Dick de Ridder, editors. Structural, Syntactic, and Statistical Pattern Recognition: Joint IAPR International Workshops, SSPR 2006 and SPR 2006, Hong Kong, China, August 17–19, 2006. Proceedings, volume 4109 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37236-9 (print), 3-540-37241-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        37241-7.

Yagi:2007:CVAa

[YKKZ07a] Yasushi Yagi, Sing Bing Kang, In So Kweon, and Hongbin Zha, editors. Computer Vision — ACCV 2007: 8th Asian Conference on Computer Vision, Tokyo, Japan, November 18–22, 2007, Proceedings, Part I, volume 4843 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76385-6 (print), 3-540-76386-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76386-4.

Yagi:2007:CVAb

[YKKZ07b] Yasushi Yagi, Sing Bing Kang, In So Kweon, and Hongbin Zha, editors. Computer Vision — ACCV 2007: 8th Asian Conference on Computer Vision, Tokyo, Japan, November 18–22, 2007, Proceedings, Part II, volume 4844 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76389-9 (print), 3-540-76390-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
76390-1.

Yu:2006:AWA

[YKL06] Jeffrey Xu Yu, Masaru Kitsuregawa, and Hong Va Leong, editors. Advances in Web-Age Information Management: 7th International Conference, WAIM 2006, Hong Kong, China, June 17–19, 2006. Proceedings, volume 4016 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-35225-2 (print), 3-540-35226-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
35226-6.

Youn:2006:UCS

[YKM06] Hee Yong Youn, Minkoo Kim, and Hiroyuki Morikawa, editors. Ubiquitous Computing Systems: Third International Symposium, UCS 2006, Seoul, Korea, October 11–13, 2006. Proceedings, volume 4239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46287-2 (print), 3-540-46289-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46289-7.

Ystad:2009:CMM

[YKMJ09] Sølvi Ystad, Richard Kronland-Martinet, and Kristoffer Jensen, editors. Computer Music Modeling and Retrieval. Genesis of Meaning in Sound and Music: 5th International Symposium, CMMR 2008 Copenhagen, Denmark, May 19–23, 2008 Revised Papers, volume 5493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02517-X (print), 3-642-02518-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
02518-1.

Yu:2011:DSAa

[YKU11a] Jeffrey Xu Yu, Myoung Ho Kim, and Rainer Unland, editors. Database Systems for Advanced Applications: 16th International Conference, DASFAA 2011, Hong Kong, China, April 22–25, 2011, Proceedings, Part I, volume 6587 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20148-2 (print), 3-642-20149-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
20149-3.

Yu:2011:DSAb

[YKU11b] Jeffrey Xu Yu, Myoung Ho Kim, and Rainer Unland, editors. Database Systems for Advanced Applications: 16th International Conference, DASFAA 2011, Hong Kong, China, April 22– 25, 2011, Proceedings, Part II, volume 6588 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-20151-2 (print), 3-642-20152-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        20152-3.

Yu:2010:UIC

[YLC^{+}10] Zhiwen Yu, Ramiro Liscano, Guanling Chen, Daqing Zhang, and Xingshe Zhou, editors. Ubiquitous Intelligence and Computing: 7th International Conference, UIC 2010, Xi’an, China, October 26–29, 2010. Proceedings, volume 6406 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16354-8 (print), 3-642-16355-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        16355-5.

Yung:2009:ISC

[YLL09] Moti Yung, Peng Liu, and Dongdai Lin, editors. Information Security and Cryptology: 4th International Conference, Inscrypt 2008, Beijing, China, December 14–17, 2008, Revised Selected Papers, volume 5487 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-01439-9 (print), 3-642-01440-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
01440-6.

Yu:2004:AWT

[YLLZ04] Jeffrey Xu Yu, Xuemin Lin, Hongjun Lu, and Yanchun Zhang, editors. Advanced Web Technologies and Applications: 6th Asia-Pacific Web Conference, APWeb 2004: Hangzhou, China, April 14–17, 2004: Proceedings, volume 3007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-21371-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .A785 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3007.
htm;   http://www.springerlink.
com/content/978-3-540-
21371-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3007;
 http://www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b96838.

Yap:2012:MBI

[YLS^{+}12] Pew-Thian Yap, Tianming Liu, Dinggang Shen, Carl-Fredrik Westin, and Li Shen, editors. Multimodal Brain Image Analysis: Second International Workshop, MBIA 2012, Held in Conjunction with MIC-CAI 2012, Nice, France, October 1–5, 2012. Proceedings, volume 7509 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33529-2 (print), 3-642-33530-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33530-3.

Yao:2007:RSK

[YLW^{+}07] JingTao Yao, Pawan Lingras, Wei-Zhi Wu, Marcin Szczuka, Nick J. Cercone, and Dominik Ślȩzak, editors. Rough Sets and Knowledge Technology: Second International Conference, RSKT 2007, Toronto, Canada, May 14–16, 2007. Proceedings, volume 4481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72457-5 (print), 3-540-72458-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72458-2.

Yeung:2006:AML

[YLWY06] Daniel S. Yeung, Zhi-Qiang Liu, Xi-Zhao Wang, and Hong Yan, editors. Advances in Machine Learning and Cybernetics: 4th International Conference, ICMLC 2005, Guangzhou, China, August 18–21, 2005, Revised Selected Papers, volume 3930 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33584-6 (print), 3-540-33585-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
33585-6.

Yonezawa:2001:MAS

[YM01] Akinori Yonezawa and Satoshi Matsuoka, editors. Metalevel architectures and separation of crosscutting concerns: Third International Conference, RE-FLECTION 2001, Kyoto, Japan, September 25–28, 2001: Proceedings, volume 2192 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42618-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2192; QA76.7 .R46 2001. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2192.htm;
 http://www.springerlink.
com/content/978-3-540-
42618-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2192.

Yang:2010:ECI

[YMHH10] Hyun Seung Yang, Rainer Malaka, Junichi Hoshino, and Jung Hyun Han, editors. Entertainment Computing — ICEC 2010: 9th International Conference, ICEC 2010, Seoul, Korea, September 8–11, 2010. Proceedings, volume 6243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15398-4 (print), 3-642-15399-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15399-0.

Yoshikawa:2010:DSA

[YMY^{+}10] Masatoshi Yoshikawa, Xiaofeng Meng, Takayuki Yumoto, Qiang Ma, Lifeng Sun, and Chiemi Watanabe, editors. Database Systems for Advanced Applications: 15th International Conference, DASFAA 2010, International Workshops: GDM, BenchmarX, MCIS, SNSMW, DIEW, UDM, Tsukuba, Japan, April 1– 4, 2010, Revised Selected Papers, volume 6193 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14588-4 (print), 3-642-14589-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14589-6.

Yakhno:2006:AIS

[YN06] Tatyana Yakhno and Erich J. Neuhold, editors. Advances in Information Systems: 4th International Conference, ADVIS 2006, Izmir, Turkey, October 18–20,2006. Proceedings, volume 4243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46291-0 (print), 3-540-46292-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46292-7.

Yahia:2008:CLT

[YNB08] Sadok Ben Yahia, Engelbert Mephu Nguifo, and Radim Belohlavek, editors. Concept Lattices and Their Applications: Fourth International Conference, CLA 2006 Tunis, Tunisia, October 30–November 1, 2006 Selected Papers, volume 4923 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78920-0 (print), 3-540-78921-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

com/content/978-3-540-
78921-5.

Yorav:2008:HSV

[Yor08] Karen Yorav, editor.

Hardware and Software: Verification and Testing: Third International Haifa Verification Conference, HVC 2007, Haifa, Israel, October 23–25, 2007. Proceedings, volume 4899 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-77964-7 (print), 3-540-77966-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77966-7.

Yu:2001:IAA

[YP01] Sheng Yu and Andrei Paun, editors. Implementation and application of automata: 5th International Conference, CIAA 2000, London, Ontario, Canada, July 24–25, 2000: Revised Papers, volume 2088 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42491-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2088; QA267 .I5787 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2088.htm;
 http://www.springerlink.
com/content/978-3-540-
42491-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2088.

Yang:2005:HPC

[YRMD05] Laurence Tianruo Yang, Omer F. Rana, Beniamino Di Martino, and Jack Dongarra, editors. High performance computing and communications: first international conference, HPCC 2005, Sorrento, Italy, September 21–23, 2005: proceedings, volume 3726 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29031-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H655 2005. URL http:

//www.springerlink.
com/content/978-3-540-
29031-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3726.

Yao:2011:RSK

[YRWS11] JingTao Yao, Sheela Ramanna, Guoyin Wang, and Zbigniew Suraj, editors. Rough Sets and Knowledge Technology: 6th International Conference, RSKT 2011, Banff, Canada, October 9–12, 2011. Proceedings, volume 6954 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24424-6 (print), 3-642-24425-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
24425-4.

Yazici:2003:CIS

[YŞ03] Adnan Yazici and Cevat Şener, editors. Computer and Information Sciences—ISCIS 2003: 18th International Symposium Antalya, Turkey, November 3–5, 2003: Proceedings, volume 2869 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20409-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2869.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        20409-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2869;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b14229.

Yoshida:2012:AICa

[YSL12] Hiroyuki Yoshida, Georgios Sakas, and Marius George Linguraru, editors. Abdominal Imaging. Computational and Clinical Applications: Third International Workshop, Held in Conjunction with MIC-CAI 2011, Toronto, ON, Canada, September 18, 2011, Revised Selected Papers, volume 7029 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28556-2 (print), 3-642-28557-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28557-8.

Yao:2010:BII

[YSP^{+}10] Yiyu Yao, Ron Sun, Tomaso Poggio, Jiming Liu, Ning Zhong, and Jimmy Huang, editors. Brain Informatics: International Conference, BI 2010, Toronto, ON, Canada, August 28– 30, 2010. Proceedings, volume 6334 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15313-5 (print), 3-642-15314-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15314-3.

Yoshiura:2006:AIC

[YSR^{+}06] Hiroshi Yoshiura, Kouichi Sakurai, Kai Rannenberg, Yuko Murayama, and Shinichi Kawamura, editors. Advances in Information and Computer Security: First International Workshop on Security, IWSEC 2006, Kyoto, Japan, October 23–24, 2006. Proceedings, volume 4266 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-47699-7 (print), 3-540-47700-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        47700-6.

Yap:1999:ACS

[YT99] Roland H. C. Yap and P. S. Thiagarajan, editors. Advances in computing science — ASIAN’99: 5th Asian Computing Science Conference, Phuket, Thailand, December 10–12, 1999: Proceedings, volume 1742 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-66856-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1742. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1742.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        66856-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1742.

Yin:2007:IDE

[YTC^{+}07] Hujun Yin, Peter Tino, Emilio Corchado, Will Byrne, and Xin Yao, editors. Intelligent Data Engineering and Automated Learning — IDEAL 2007: 8th International Conference, Birmingham, UK, December 16–19, 2007. Proceedings, volume 4881 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77225-1 (print), 3-540-77226-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77226-2.

Ytrehus:2006:CCI

[Ytr06] Øyvind Ytrehus, editor. Coding and Cryptography: International Workshop, WCC 2005, Bergen, Norway, March 14–18, 2005. Revised Selected Papers, volume 3969 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

URL http://www.springerlink.
com/content/978-3-540-
35481-9.

Yu:2002:HDI

[Yu02] Cui Yu. High-Dimensional Indexing: Transformational Approaches to High-Dimensional Range and Similarity Searches, volume 2341 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44199-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 150 pp. LCCN QA76.9.D3 Y83 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2341.
htm;   http://www.springerlink.
com/content/978-3-540-
44199-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2341.

Yung:2002:ACC

[Yun02] Moti Yung, editor. Advances in cryptology — CRYPTO 2002: 22nd Annual International Cryptology Conference, Santa Barbara, California, USA, August 18–22, 2002: Proceedings, volume 2442 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44050-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 C79 2002. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t2442.htm;
http://www.springerlink.
com/content/978-3-540-
44050-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2442.

Yang:2006:PTA

[YW06] Qiang Yang and Geoff Webb, editors. PRI-CAI 2006: Trends in Artificial Intelligence: 9th Pacific Rim International Conference on Artificial Intelligence Guilin, China, August 7–11, 2006 Proceedings, volume 4099 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36667-9 (print), 3-540-36668-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36668-3.

Yin:2004:ANNa

[YWG04a] Fuliang Yin, Jun Wang, and Chengan Guo, editors. Advances in Neural Networks—ISNN 2004: International Symposium on Neural Networks, Dalian, China, August 19–21, 2004: Proceedings, Part I, volume 3173 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22841-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.87 .I5845 2004; TA1650 .D35 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3173.htm;
 http://www.springerlink.
com/content/978-3-540-
22841-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3173;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99834.

Yin:2004:ANNb

[YWG04b] Fuliang Yin, Jun Wang, and Chengan Guo, editors. Advances in Neural Networks—ISNN 2004: International Symposium on Neural Networks, Dalian, China, August 19–21, 2004: Proceedings, Part II, volume 3174 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22843-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3174.htm;
http://www.springerlink.
com/content/978-3-540-
22843-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3174;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99834.

Yin:2011:IDE

[YWRS11] Hujun Yin, Wenjia Wang, and Victor Rayward-Smith, editors. Intelligent Data Engineering and Automated Learning

— IDEAL 2011: 12th International Conference, Norwich, UK, September 7–9, 2011. Proceedings, volume 6936 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23877-7 (print), 3-642-23878-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23878-9.

Yew:2004:ACS

[YX04] Pen-Chung Yew and Jingling Xue, editors. Advances in Computer Systems Architecture: 9th Asia-Pacific Conference, ACSAC 2004, Beijing, China, September 7– 9, 2004: Proceedings, volume 3189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23003-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A73 A28; QA76 A1 L43 3189. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3189.htm;
 http://www.springerlink.
com/content/978-3-540-
23003-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3189;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100354.

Yuan:2001:IAS

[YY01] Soe-Tsyr Yuan and Makoto Yokoo, editors. Intelligent agents: specification, modeling, and applications: 4th Pacific Rim International Workshop on Multi-Agents, PRIMA 2001, Taipei, Taiwan, July 28–29, 2001: Proceedings, volume 2132 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42434-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2132; QA76.76.I58 P75 2001.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2132.
htm;   http://www.springerlink.
com/content/978-3-540-
42434-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2132.

Youm:2009:ISA

[YY09] Heung Youl Youm and Moti Yung, editors. Information Security Applications: 10th International Workshop, WISA 2009, Busan, Korea, August 25–27, 2009, Revised Selected Papers, volume 5932 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-10837-7 (print), 3-642-10838-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
10838-9.

Yu:2012:DSA

[YYH^{+}12] Hwanjo Yu, Ge Yu, Wynne Hsu, Yang-Sae Moon, Rainer Unland, and Jaesoo Yoo, editors. Database Systems for Advanced Applications: 17th International Conference, DAS-FAA 2012, International Workshops: FlashDB, ITEMS, SNSM, SIM 3, DQDI, Busan, South Korea, April 15–19, 2012. Proceedings, volume 7240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29022-1 (print), 3-642-29023-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29023-7.

Yang:2009:PPM

[YYI^{+}09] Jung-Jin Yang, Makoto Yokoo, Takayuki Ito, Zhi Jin, and Paul Scerri, editors. Principles of Practice in Multi-Agent Systems: 12th International Conference, PRIMA 2009, Nagoya, Japan, December 14– 16, 2009. Proceedings, volume 5925 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-11160-2 (print), 3-642-11161-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11161-7.

Yao:2012:RSC

[YYS^{+}12] JingTao Yao, Yan Yang, Roman Słowiński, Salvatore Greco, Huaxiong Li, Sushmita Mitra, and Lech Polkowski, editors. Rough Sets and Current Trends in Computing: 8th International Conference, RSCTC 2012, Chengdu, China, August 17–20, 2012.Proceedings, volume 7413 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32114-3 (print), 3-642-32115-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
32115-3.

Yang:2007:ISI

[YZC^{+}07] Christopher C. Yang, Daniel Zeng, Michael Chau, Kuiyu Chang, Qing Yang, Xueqi Cheng, Jue Wang, Fei-Yue Wang, and Hsinchun Chen, editors. Intelligence and Security Informatics: Pacific Asia Workshop, PAISI 2007, Chengdu, China, April 11–12, 2007. Proceedings, volume 4430 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71548-7 (print), 3-540-71549-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71549-8.

Yuille:2007:EMM

[YZCW07] Alan L. Yuille, Song-Chun Zhu, Daniel Cremers, and Yongtian Wang, editors. Energy Minimization Methods in Computer Vision and Pattern Recognition: 6th International Conference, EMMCVPR 2007, Ezhou, China, August 27–29, 2007. Proceedings, volume 4679 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74195-X (print), 3-540-74198-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
74198-5.

Yang:2005:ESS

[YZZ^{+}05] Laurence Tianruo Yang, Xingshe Zhou, Wei Zhao, Zhaohui Wu, Yian Zhu, and Man Lin, editors. Embedded software and systems: second international conference, ICESS 2005, Xi’an, China, December 16–18, 2005: proceedings, volume 3820 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30881-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.E42 I33 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30881-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3820.

Zamboni:2008:DIM

[Zam08] Diego Zamboni, editor.

Detection of Intrusions and Malware, and Vulnerability Assessment: 5th International Conference, DIMVA 2008, Paris, France, July 10– 11, 2008. Proceedings, volume 5137 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-70541-4 (print), 3-540-70542-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
70542-0.

Zaniolo:2000:ADT

[Zan00] Carlo Zaniolo, editor. Advances in database technology — EDBT 2000: 7th International Conference on Extending Database Technology, Konstanz, Germany, March 27–31, 2000: Proceedings, volume 1777 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67227-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55844 2000; QA267.A1 L43 no.1777. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1777.htm;
 http://www.springerlink.
com/content/978-3-540-
67227-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1777.

Zlatuska:1998:MFC

[ZBG98] Jiří Zlatuška, Lubos Brim, and Jozef Gruska, editors. Mathematical foundations of computer science 1998: 23rd International Symposium, MFCS ’98, Brno, Czech Republic, August 24– 28, 1998: Proceedings, volume 1450 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64827-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 I58 1998.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1450.
htm;   http://www.springerlink.
com/content/978-3-540-
64827-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1450.

Ziemke:2012:AAI

[ZBH12] Tom Ziemke, Christian Balkenius, and John Hallam, editors. From Animals to Animats 12: 12th International Conference on Simulation of Adaptive Behavior, SAB 2012, Odense, Denmark, August 27–30, 2012. Proceedings, volume 7426 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33092-4 (print), 3-642-33093-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33093-3.

Zaphiris:2012:TPD

[ZBRL12] Panayiotis Zaphiris, George Buchanan, Edie Rasmussen, and Fernando Loizides, editors. Theory and Practice of Digital Libraries: Second International Conference, TPDL 2012, Paphos, Cyprus, September 23– 27, 2012. Proceedings, volume 7489 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33289-7 (print), 3-642-33290-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        33290-6.

Zagler:1994:CHP

[ZBW94] Wolfgang Zagler, Geoffrey Busby, and Roland R. Wagner, editors. Computers for handicapped persons: 4th International Conference, IC-CHP ’94, Vienna, Austria, September 14–16, 1994: proceedings, volume 860 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58476-5 (Berlin), 0-387-58476-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HV1569.5 .I56 1994. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0860.htm;
 http://www.springerlink.
com/content/978-3-540-
58476-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=860.

Papers presented at the Fourth International Conference on Computers for Handicapped Persons held in Vienna, Austria, in 1994.

Ziviani:2007:SPI

[ZBY07] Nivio Ziviani and Ricardo Baeza-Yates, editors. String Processing and Information Retrieval: 14th International Symposium, SPIRE 2007 Santiago, Chile, October 29–31, 2007 Proceedings, volume 4726 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75529-2 (print), 3-540-75530-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        75530-2.

Zeadally:2010:FMN

[ZCCL10] Sherali Zeadally, Eduardo Cerqueira, Marília Curado, and Mikołaj Leszczuk, editors. Future Multimedia Networking: Third International Workshop, FMN 2010, Kraków, Poland, June 17–18, 2010. Proceedings, volume 6157 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13788-1 (print), 3-642-13789-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        13789-1.

Zhang:2010:HPC

[ZCDT10] Wu Zhang, Zhangxin Chen, Craig C. Douglas, and Weiqin Tong, editors. High Performance Computing and Applications: Second International Conference, HPCA 2009, Shanghai, China, August 10–12, 2009, Revised Selected Papers, volume 5938 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-11841-0 (print), 3-642-11842-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
11842-5.

Zhong:2011:AMT

[ZCGH11] Ning Zhong, Vic Callaghan, Ali A. Ghorbani, and Bin Hu, editors. Active Media Technology: 7th International Conference, AMT 2011, Lanzhou, China, September 7–9, 2011. Proceedings, volume 6890 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23619-7 (print), 3-642-23620-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
23620-4.

Zeng:2008:BBI

[ZCRL08] Daniel Zeng, Hsinchun Chen, Henry Rolka, and Bill Lober, editors. Biosurveillance and Biosecurity: International Workshop, BioSecure 2008, Raleigh, NC, USA, December 2, 2008. Proceedings, volume 5354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89745-3 (print), 3-540-89746-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        89746-0.

Zaverucha:2008:AAI

[ZdC08] Gerson Zaverucha and Augusto Loureiro da Costa, editors. Advances in Artificial Intelligence — SBIA 2008: 19th Brazilian Symposium on Artificial Intelligence Savador, Brazil, October 26–30, 2008. Proceedings, volume 5249 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-88189-1 (print), 3-540-88190-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
88190-2.

Zannier:2004:EPA

[ZEL04] Carmen Zannier, Hakan Erdogmus, and Lowell Lindstrom, editors. Extreme programming and agile methods—XP/Agile Universe 2004: 4th Conference on Extreme Programming and Agile Methods, Calgary, Canada, August 15– 18, 2004: proceedings, volume 3134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22839-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 X63 2004. URL http:

//www.springerlink.
com/content/978-3-540-
22839-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3134;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99820.

Zhuge:2005:GCC

[ZF05] Hai Zhuge and Geoffrey C. Fox, editors. Grid and Cooperative Computing: GCC 2005: 4th International Conference, Beijing, China, November 30–December 3, 2005: Proceedings, volume 3795 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-30510-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C58 G34 2005. URL http:

//www.springerlink.
com/content/978-3-540-
30510-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3795.

Zhao:2003:IPS

[ZG03] Feng Zhao and Leonidas J. Guibas, editors. Information processing in sensor networks: Second International Workshop, IPSN 2003, Palo Alto, CA, USA, April 22–23, 2003: Proceedings, volume 2634 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-02111-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7872.D48 I64 2003. URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t2634.htm;
http://www.springerlink.
com/content/978-3-540-
02111-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2634.

Also available via the World Wide Web.

Zeng:2007:ISI

[ZGK^{+}07] Daniel Zeng, Ivan Gotham, Ken Komatsu, Cecil Lynch, Mark Thurmond, David Madigan, Bill Lober, James Kvach, and Hsinchun Chen, editors. Intelligence and Security Informatics: Biosurveillance: Second NSF Workshop, Bio-Surveillance 2007, New Brunswick, NJ, USA, May 22, 2007. Proceedings, volume 4506 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-72607-1 (print), 3-540-72608-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
72608-1.

Zhao:2005:AMF

[ZGT05] Wenyi Zhao, Shaogang Gong, and Xiaoou Tang, editors. Analysis and modelling of faces and gestures: second international workshop, AMFG 2005, Beijing, China, October 16, 2005. proceedings, volume 3723 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29229-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
29229-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3723.

Zhang:2004:PTA

[ZGY04] Chengqi Zhang, Hans W. Guesgen, and Wai K. Yeap, editors. PRICAI 2004: Trends in Artificial Intelligence: 8th Pacific Rim International Conference on Artificial Intelligence, Auckland, New Zealand, August 9–13, 2004. Proceedings, volume 3157 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22817-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//link.springer-ny.

com/link/service/series/
0558/tocs/t3157.htm;
 http://www.springerlink.
com/content/978-3-540-
22817-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3157;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99563.

Zaki:2000:LSP

[ZH00] Mohammed J. Zaki and Ching-Tien Ho, editors. Large-Scale Parallel Data Mining, volume 1759 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67194-3 (alk. paper). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343 L37 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1759.htm;
 http://www.springerlink.
com/content/978-3-540-
67194-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1759.

Based on the workshop on Large-Scale Parallel KDD Systems, which was held in conjunction with the 5th ACM SIGKDD International Conference on Knowledge Discovery and Data Mining, on August 15th, 1999, San Diego, California.

Zhang:2007:MBF

[Zha07] David Zhang, editor. Medical Biometrics: First International Conference, ICMB 2008, Hong Kong, China, January 4–5, 2008. Proceedings, volume 4901 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77410-6 (print), 3-540-77413-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77413-6.

Zheng:2002:ACA

[Zhe02] Yuliang Zheng, editor. Advances in Cryptology—ASIACRYPT 2002: 8th International Conference on the Theory and Application of Cryptology and Information Security, Queenstown, New Zealand, December 1–5, 2002. Proceedings, volume 2501 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-00171-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I555 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2501.htm;
 http://www.springerlink.
com/content/978-3-540-
00171-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2501.

Also available via the World Wide Web.

Zhang:2004:CIS

[ZHF04] Jun Zhang, Ji-Huan He, and Yuxi Fu, editors. Computational and information science: first international symposium, CIS 2004, Shanghai, China, December 16–18, 2004: proceedings, volume 3314 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-24127-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I5843 2004. URL

http://springerlink.
metapress.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3314;
 http://www.springerlink.
com/content/978-3-540-
24127-0;   http://www.

                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3314;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b104566.

Zhang:2012:ABI

[ZHLW12] Huaguang Zhang, Amir Hussain, Derong Liu, and Zhanshan Wang, editors. Advances in Brain Inspired Cognitive Systems: 5th International Conference, BICS 2012, Shenyang, China, July 11–14, 2012. Proceedings, volume 7366 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31560-7 (print), 3-642-31561-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31561-9.

Zhong:2001:WIR

[Zho01] Ning Zhong, editor. Web intelligence: research and development: first Asia-Pacific Conference, WI 2001, Maebashi City, Japan, October 23–26, 2001: Proceedings, volume 2198 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42730-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .W49 2001; QA267.A1 L43 no.2198. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2198.htm;
 http://www.springerlink.
com/content/978-3-540-
42730-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2198.

Zsok:2012:CEF

[ZHP12] Viktória Zsók, Zoltán Horváth, and Rinus Plasmeijer, editors. Central European Functional Programming School: 4th Summer School, CEFP 2011, Budapest, Hungary, June 14–24, 2011, Revised Selected Papers, volume 7241 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32095-3 (print), 3-642-32096-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        32096-5.

Zilli:1987:MMS

[Zil87] Marisa Venturini Zilli, editor. Mathematical models for the semantics of parallelism: Advanced School, Rome, Italy, September 24– October 1, 1986: proceedings, volume 280 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1987. CO-DEN LNCSD9. ISBN 3-540-18419-8 (Berlin), 0-387-18419-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .A3331 1986. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0280.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        18419-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=280.

“Presented at the Advanced School on Mathematical Models for the Semantics of Parallelism held in Rome at the Istituto per le Applicazioni del Calcolo ‘Mauro Picone”’–Pref.

Zima:1992:PCF

[Zim92] Hans Zima, editor. Parallel computation: First International ACPC Conference, Salzburg, Austria, September 30– October 2, 1991: proceedings, volume 591 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 3-540-55437-8 (Berlin), 0-387-55437-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5 1991.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t0591.
htm;   http://www.springerlink.
com/content/978-3-540-
55437-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=591.

Zippel:1992:CAP

[Zip92] R. E. Zippel, editor.

Computer algebra and parallelism: second international workshop, Ithaca, USA, May 9– 11, 1990: proceedings, volume 584 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55328-2 (Berlin), 0-387-55328-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 C6495 1992. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0584.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        55328-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=584.

Zitzler:2001:EMC

[Zit01] Eckart Zitzler, editor. Evolutionary multi-criterion optimization: First International Conference, EMO 2001, Zürich, Switzerland, March 7– 9, 2001: Proceedings, volume 1993 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-41745-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T57.95 .E43 2001 Stacks SEMSTAX-STAX; QA267.A1 L43 no.1993; T57.95 .I57 2001. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1993.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41745-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1993.

Zha:2010:CVAa

[ZiTM10a] Hongbin Zha, Rin ichiro Taniguchi, and Stephen Maybank, editors. Computer Vision — ACCV 2009: 9th Asian Conference on Computer Vision, Xi’an, September 23–27, 2009, Revised Selected Papers, Part I, volume 5994 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12306-6 (print), 3-642-12307-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
12307-8.

Zha:2010:CVAb

[ZiTM10b] Hongbin Zha, Rin ichiro Taniguchi, and Stephen Maybank, editors. Computer Vision — ACCV 2009: 9th Asian Conference on Computer Vision, Xi’an, September 23–27, 2009, Revised Selected Papers, Part II, volume 5995 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12303-1 (print), 3-642-12304-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12304-7.

Zha:2010:CVAc

[ZiTM10c] Hongbin Zha, Rin ichiro Taniguchi, and Stephen Maybank, editors. Computer Vision — ACCV 2009: 9th Asian Conference on Computer Vision, Xi’an, September 23–27, 2009, Revised Selected Papers, Part III, volume 5996 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-12296-5 (print), 3-642-12297-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        12297-2.

Zhang:2004:BAF

[ZJ04a] David Zhang and Anil K. Jain, editors. Biometric Authentication: First International Conference, ICBA 2004, Hong Kong, China, July 15–17, 2004, Proceedings, volume 3072 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22146-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 A1 L43 3072 (LC). URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3072.htm;
 http://www.springerlink.
com/content/978-3-540-
22146-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3072;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98225.

Zhang:2004:WSE

[ZJ04b] Liang-Jie (LJ) Zhang and Mario Jeckle, editors. Web Services: European Conference, ECOWS 2004, Erfurt, Germany, September 27– 30, 2004. Proceedings, volume 3250 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23202-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.88813 .E26 2004.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t3250.
htm;   http://www.springerlink.
com/content/978-3-540-
23202-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3250;

                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100919.

Zhang:2005:AAA

[ZJ05] Shichao Zhang and Raymond A. (Raymond Austin) Jarvis, editors. AI 2005: advances in artificial intelligence: 18th Australian Joint Conference on Artificial Intelligence, Sydney, Australia, December 5–9, 2005: proceedings, volume 3809 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-30462-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A96 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        30462-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3809.

Zhang:2006:ABI

[ZJ06] David Zhang and Anil K. Jain, editors. Advances in biometrics: international conference, ICB 2006, Hong Kong, China, January 5–7, 2006: proceedings, volume 3832 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31111-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.B56 I57 2006.

URL http://www.springerlink.
com/content/978-3-540-
31111-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3832.

Zheng:2006:AMV

[ZJL06] Nanning Zheng, Xiaoyi Jiang, and Xuguang Lan, editors. Advances in Machine Vision, Image Processing, and Pattern Analysis: International Workshop on Intelligent Computing in Pattern Analysis/Synthesis, IWICPAS 2006 Xi’an, China, August 26–27, 2006 Proceedings, volume 4153 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-37597-X (print), 3-540-37598-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
37598-2.

Zima:2002:HPC

[ZJS^{+}02] Hans P. Zima, Kazuki Joe, Mitsuhisa Sato, Yoshiko Seo, and Masaaki Shimasaki, editors. High Performance Computing: 4th International Symposium, ISHPC 2002, Kansai Science City, Japan, May 15–17, 2002. Proceedings, volume 2327 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43674-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I585 2002. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2327.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        43674-4;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2327.

Zhou:2003:APP

[ZJXC03] Xingming Zhou, Stefan Jähnichen, Ming Xu, and Jiannong Cao, editors. Advanced Parallel Processing Technologies: 5th International Workshop, APPT 2003, Xiamen, China, September 17–19, 2003: Proceedings, volume 2834 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20054-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2834.htm;
 http://www.springerlink.
com/content/978-3-540-
20054-3;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2834;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b13582.

Zamboni:2006:RAI

[ZK06] Diego Zamboni and Christopher Kruegel, editors. Recent Advances in Intrusion Detection: 9th International Symposium, RAID 2006 Hamburg, Germany, September 20–22, 2006 Proceedings, volume 4219 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-39723-X (print), 3-540-39725-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-540-
                                        39725-0.

Zighed:2000:PDM

[ZKZ00] Djamel A. Zighed, J. Komorowski, and Jan M. Zytkow, editors. Principles of Data Mining and Knowledge Discovery: 4th European Conference, PKDD 2000, Lyon, France, September 13–16, 2000. Proceedings, volume 1910 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-41066-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1910; QA76.9.D3 P53 2000. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1910.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        41066-9;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1910.

Zhang:1996:DAI

[ZL96] Chengqi Zhang and Dickson Lukose, editors. Distributed artificial intelligence: architecture and modelling: First Australian Workshop on DAI, Canberra, ACT, Australia, November 13, 1995: proceedings, volume 1087 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CO-DEN LNCSD9. ISBN 3-540-61314-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q337 .A88 1995.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1087.
htm;   http://www.springerlink.
com/content/978-3-540-
61314-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1087.

Zhang:1997:MAS

[ZL97] Chengqi Zhang and Dickson. Lukose, editors. Multi-agent systems: methodologies and applications: second Australian Workshop on Distributed Artificial Intelligence, Cairns, QLD, Australia, August 27, 1996: Selected Papers, volume 1286 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63412-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5A97 1996. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1286.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        63412-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1286.

Zhang:1998:MAS

[ZL98] Chengqi Zhang and Dickson Lukose, editors. Multi-agent systems: theories, languages, and applications: 4th Australian Workshop on Distributed Artificial Intelligence, Brisbane, Qld., Australia, July 13, 1998: Proceedings, volume 1544 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65477-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58A97 1998. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1544.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65477-3;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1544.

Zelezny:2008:ILP

[ŽL08] Filip Železný and Nada Lavrač, editors. Inductive Logic Programming: 18th International Conference, ILP 2008 Prague, Czech Republic, September 10–12, 2008 Proceedings, volume 5194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-85927-6 (print), 3-540-85928-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
85928-4.

Zhou:2005:ISI

[ZLDB05] Jianying Zhou, Javier López, Robert Deng, and Feng Bao, editors. Information security: 8th international conference, ISC 2005, Singapore, September 20–23, 2005: proceedings, volume 3650 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-29001-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I85 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        29001-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3650.

Zhang:2005:CVB

[ZLJ05] Changshui Zhang, Yanxi Liu, and Tianxi Jiang, editors. Computer vision for biomedical image applications: first international workshop, CVBIA 2005, Beijing, China, October 21, 2005: proceedings, volume 3765 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-29411-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN R859.7.C67 C83 2005.

                                        URL http://www.springerlink.
                                        com/content/978-3-540-
                                        29411-5;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3765.

Zhang:2010:ANNa

[ZLK10a] Liqing Zhang, Bao-Liang Lu, and James Kwok, editors. Advances in Neural Networks — ISNN 2010: 7th International Symposium on Neural Networks, ISNN 2010, Shanghai, China, June 6–9, 2010, Proceedings, Part I, volume 6063 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13277-4 (print), 3-642-13278-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13278-0.

Zhang:2010:ANNb

[ZLK10b] Liqing Zhang, Bao-Liang Lu, and James Kwok, editors. Advances in Neural Networks — ISNN 2010: 7th International Symposium on Neural Networks, ISNN 2010, Shanghai, China, June 6–9, 2010, Proceedings, Part II, volume 6064 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13317-7 (print), 3-642-13318-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13318-3.

Zhong:2009:BII

[ZLLC09] Ning Zhong, Kuncheng Li, Shengfu Lu, and Lin Chen, editors. Brain Informatics: International Conference, BI 2009 Beijing, China, October 22–24, 2009 Proceedings, volume 5819 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-04953-2 (print), 3-642-04954-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
04954-5.

Zhou:2006:FWR

[ZLS^{+}06] Xiaofang Zhou, Jianzhong Li, Heng Tao Shen, Masaru Kitsuregawa, and Yanchun Zhang, editors. Frontiers of WWW research and development–APWeb 2006: 8th Asia-Pacific Web Conference, Harbin, China, January 16–18, 2006: proceedings, volume 3841 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31142-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .A785 2006.

URL http://www.springerlink.
com/content/978-3-540-
31142-3;   http://www.
springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=3841.

Zhong:2007:WIM

[ZLY^{+}07a] Ning Zhong, Jiming Liu, Yiyu Yao, Jinglong Wu, Shengfu Lu, and Kuncheng Li, editors. Web Intelligence Meets Brain Informatics: First WICI International Workshop, WImBI 2006, Beijing, China, December 15–16, 2006, Revised Selected and Invited Papers, volume 4845 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77027-5 (print), 3-540-77028-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
77028-2.

Zhou:2007:AKD

[ZLY07b] Zhi-Hua Zhou, Hang Li, and Qiang Yang, editors. Advances in Knowledge Discovery and Data Mining: 11th Pacific-Asia Conference, PAKDD 2007, Nanjing, China, May 22–25, 2007. Proceedings, volume 4426 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-71700-5 (print), 3-540-71701-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
71701-0.

Zhou:2003:AWB

[ZNCF03] Wanlei Zhou, Paul Nicholson, Brian Corbitt, and Joseph Fong, editors. Advances in Web-Based Learning—ICWL 2003: Second International Conference, Melbourne, Australia, August 18–20, 2003: Proceedings, volume 2783 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-40772-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN LB1044.87 .I39 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2783.
htm;   http://www.springerlink.
com/content/978-3-540-
40772-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2783;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12012.

Zhang:2010:PTA

[ZO10] Byoung-Tak Zhang and Mehmet A. Orgun, editors. PRICAI 2010: Trends in Artificial Intelligence: 11th Pacific Rim International Conference on Artificial Intelligence, Daegu, Korea, August 30–September 2, 2010. Proceedings, volume 6230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-15245-7 (print), 3-642-15246-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
15246-7.

Zhang:2007:MAH

[ZOCJ07] Hongke Zhang, Stephan Olariu, Jiannong Cao, and David B. Johnson, editors. Mobile Ad-Hoc and Sensor Networks: Third International Conference, MSN 2007 Beijing, China, December 12–14, 2007 Proceedings, volume 4864 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-77023-2 (print), 3-540-77024-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        77024-4.

Zhou:2005:DSA

[ZOM05] Lizhu Zhou, Beng Chin Ooi, and Xiaofeng Meng, editors. Database systems for advanced applications: 10th international conference, DASFAA 2005, Beijing, China, April 17– 20, 2005: Proceedings, volume 3453 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-25334-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5533 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        25334-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3453;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b107189.

Zha:2006:ITS

[ZPT^{+}06] Hongbin Zha, Zhigeng Pan, Hal Thwaites, Alonzo C. Addison, and Maurizio Forte, editors. Interactive Technologies and Sociotechnical Systems: 12th International Conference, VSMM 2006, Xi’an, China, October 18–20, 2006. Proceedings, volume 4270 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46304-6 (print), 3-540-46305-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
46305-4.

Zhang:2009:UIC

[ZPTI09] Daqing Zhang, Marius Portmann, Ah-Hwee Tan, and Jadwiga Indulska, editors. Ubiquitous Intelligence and Computing: 6th International Conference, UIC 2009, Brisbane, Australia, July 7–9, 2009. Proceedings, volume 5585 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-02829-2 (print), 3-642-02830-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        02830-4.

Zytkow:1998:PDM

[ZQ98] Jan M. Zytkow and Mohamed Quafafou, editors. Principles of data mining and knowledge discovery: Second European Symposium, PKDD ’98, Nantes, France, September 23–26, 1998: Proceedings, volume 1510 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. CO-DEN LNCSD9. ISBN 3-540-65068-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA261.A1 L43 no.1510. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t1510.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        65068-3;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=1510.

Zytkow:1999:PDM

[ŻR99] Jan M. Żytkow and Jan Rauch, editors. Principles of data mining and knowledge discovery: Third European Conference, PKDD’99, Prague, Czech Republic, September 15–18, 1999: Proceedings, volume 1704 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66490-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 P53 1999; QA267.A1 L43 no.1704. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1704.htm;
 http://www.springerlink.
com/content/978-3-540-
66490-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1704.

Zaman:2009:VIB

[ZRP^{+}09] Halimah Badioze Zaman, Peter Robinson, Maria Petrou, Patrick Olivier, Heiko Schröder, and Timothy K. Shih, editors. Visual Informatics: Bridging Research and Practice: First International Visual Informatics Conference, IVIC 2009 Kuala Lumpur, Malaysia, November 11– 13, 2009 Proceedings, volume 5857 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05035-2 (print), 3-642-05036-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        05036-7.

Zaman:2011:VISa

[ZRP^{+}11a] Halimah Badioze Zaman, Peter Robinson, Maria Petrou, Patrick Olivier, Timothy K. Shih, Sergio Velastin, and Ingela Nyström, editors. Visual Informatics: Sustaining Research and Innovations: Second International Visual Informatics Conference, IVIC 2011, Selangor, Malaysia, November 9– 11, 2011, Proceedings, Part I, volume 7066 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25190-0 (print), 3-642-25191-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25191-7.

Zaman:2011:VISb

[ZRP^{+}11b] Halimah Badioze Zaman, Peter Robinson, Maria Petrou, Patrick Olivier, Timothy K. Shih, Sergio Velastin, and Ingela Nyström, editors. Visual Informatics: Sustaining Research and Innovations: Second International Visual Informatics Conference, IVIC 2011, Selangor, Malaysia, November 9– 11, 2011, Proceedings, Part II, volume 7067 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25199-4 (print), 3-642-25200-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
25200-6.

Zhong:2003:FIS

[ZRTS03] Ning Zhong, Zbigniew W.

Raś, Shusaku Tsumoto, and Einoshin Suzuki, editors. Foundations of Intelligent Systems: 14th International Symposium, Maebashi City, Japan, October 28–31, 2003: Proceedings, volume 2871 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20256-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t2871.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20256-1;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2871;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b14019.

Zhang:2000:DAI

[ZS00] Chengqi Zhang and Von-Wun Soo, editors. Design and Applications of Intelligent Agents: Third Pacific Rim International Workshop on Multi-Agents, PRIMA 2000, Melbourne, Australia, August 28–29, 2000. Proceedings, volume 1881 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67911-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58 P75 2000.

                                        URL http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t1881.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        67911-0;   http://www.
                                        springerlink.com/openurl.

asp?genre=issue&issn=
0302-9743&volume=1881.

Zimnyi:2005:JDS

[ZS05a] Esteban Zimńyi and Stefano Spaccapietra, editors. Journal on data semantics. III, volume 3534 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26225-3. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 212 pp. LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
26225-1;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3534.

Zucker:2005:ARA

[ZS05b] Jean-Daniel Zucker and Lorenza Saitta, editors. Abstraction, reformulation and approximation: 6th international symposium, SARA 2005, Airth Castle, Scotland, UK, July 26–29, 2005; proceedings, volume 3607 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-27872-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .S27 2005. URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        27872-6;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3607.

Zhang:2007:KSE

[ZS07] Zili Zhang and Jörg Siekmann, editors. Knowledge Science, Engineering and Management: Second International Conference, KSEM 2007, Melbourne, Australia, November 28– 30, 2007. Proceedings, volume 4798 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-76718-5 (print), 3-540-76719-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        76719-0.

Zhang:2010:MBS

[ZS10] David Zhang and Milan Sonka, editors. Medical Biometrics: Second International Conference, ICMB 2010, Hong Kong, China, June 28–30, 2010. Proceedings, volume 6165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13922-1 (print), 3-642-13923-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13923-9.

Zaiane:2003:MMC

[ZSD03] Osmar R. Zaïane, Simeon J.

Simoff, and Chabane Djeraba, editors. Mining Multimedia and Complex Data: KDD workshop MDM/KDD 2002, PAKDD workshop KDMCD 2002: Revised Papers, volume 2797 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20305-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 K3588 2002.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2797.
htm;   http://www.springerlink.
com/content/978-3-540-
20305-6;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2797;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b12031.

Zhang:2009:AWM

[ZSM^{+}09] Haizheng Zhang, Myra Spiliopoulou, Bamshad Mobasher, C. Lee Giles, Andrew McCallum, Olfa Nasraoui, Jaideep Srivastava, and John Yen, editors. Advances in Web Mining and Web Usage Analysis: 9th International Workshop on Knowledge Discovery on the Web, WebKDD 2007, and 1st International Workshop on Social Networks Analysis, SNA-KDD 2007, San Jose, CA, USA, August 12– 15, 2007. Revised Papers, volume 5439 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00527-6 (print), 3-642-00528-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        00528-2.

Zhong:1999:NDR

[ZSO99] Ning Zhong, Andrzej Skowron, and Setsuo Ohsuga, editors. New directions in rough sets, data mining, and granularsoft computing: 7th International Workshop, RSFDGrC’99, Yamaguchi, Japan, November 9– 11, 1999: Proceedings, volume 1711 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66645-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .R494 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1711.htm;
 http://www.springerlink.
com/content/978-3-540-
66645-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1711.

Zhou:2004:WIS

[ZSP^{+}04] Xiaofang Zhou, Stanley Su, Mike P. Papazoglou, Maria E. Orlowska, and Keith G. Jeffery, editors. Web Information Systems—WISE 2004: 5th International Conference on Web Information Systems Engineering, Brisbane, Australia, November 22–24, 2004: Proceedings, volume 3306 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23894-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA168 .I5265 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3306.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        23894-2;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3306;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b103344.

Zseby:2010:FIF

[ZSP10] Tanja Zseby, Reijo Savola, and Marco Pistore, editors. Future Internet

— FIS 2009: Second Future Internet Symposium, FIS 2009, Berlin, Germany, September 1– 3, 2009, volume 6152 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14955-3 (print), 3-642-14956-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        14956-6.

Zaiane:2003:WMW

[ZSSM03] Osmar R. Zaïane, Jaideep Srivastava, Myra Spiliopoulou, and Brij Masand, editors. WEBKDD 2002— Mining Web Data for Discovering Usage Patterns and Profiles: 4th International Workshop Edmonton, Canada, July 23, 2002 Revised Papers, volume 2703 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-20304-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4235 .W435 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2703.htm;
http://www.springerlink.
com/content/978-3-540-
20304-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2703;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b11784.

Zhou:2006:EDE

[ZSY^{+}06] Xiaobo Zhou, Oleg Sokolsky, Lu Yan, Eun-Sun Jung, Zili Shao, Yi Mu, Dong Chun Lee, Dae Young Kim, Young-Sik Jeong, and Cheng-Zhong Xu, editors. Emerging Directions in Embedded and Ubiquitous Computing: EUC 2006 Workshops: NCUS, SecUbiq, USN, TRUST, ESO, and MSA, Seoul, Korea, August 1–4, 2006. Proceedings, volume 4097 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-36850-7 (print), 3-540-36851-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
36851-9.

Zimmermann:2004:ASM

[ZT04] Wolf Zimmermann and Bernhard Thalheim, editors. Abstract State Machines 2004. Advances in Theory and Practice: 11th International Workshop, ASM 2004, Lutherstadt Wittenberg, Germany, May 24–28, 2004: Proceedings, volume 3052 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-22094-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267 .A. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3052.htm;
 http://www.springerlink.
com/content/978-3-540-
22094-7;   http://www.

springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=3052;
http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b98118.

Zanzotto:2012:BII

[ZTTY12] Fabio Massimo Zanzotto, Shusaku Tsumoto, Niels Taatgen, and Yiyu Yao, editors. Brain Informatics: International Conference, BI 2012, Macau, China, December 4–7, 2012. Proceedings, volume 7670 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35138-7 (print), 3-642-35139-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
35139-6.

Zhang:2005:WTR

[ZTY^{+}05] Yanchun Zhang, Katsumi Tanaka, Jeffrey Xu Yu, Shan Wang, and Minglu Li, editors. Web technologies research and development, APWeb 2005: 7th Asia-Pacific Web Conference, Shanghai, China, March 29– April 1, 2005: Proceedings, volume 3399 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CO-DEN LNCSD9. ISBN 3-540-25207-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .A785 2005.

URL http://www.springerlink.
com/content/978-3-540-
25207-8;   http://www.
springerlink.com/link.
asp?id=pcrnvhf4(version)
;   http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3399;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b106936.

Zuck:2003:VMC

[Zuc03] Lenore D. Zuck, editor. Verification, model checking, and abstract interpretation: 4th International Conference, VMCAI 2003, New York, NY, USA, January 9– 11, 2003: Proceedings, volume 2575 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-00348-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 V53 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2575.htm;
http://www.springerlink.
com/content/978-3-540-
00348-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2575.

Zuidweg:1999:ISN

[Zui99] Han Zuidweg, editor. Intelligence in services and networks: paving the way for an open service market: 6th International Conference on Intelligence and Services in Networks, IS and N’99, Barcelona, Spain, April 27–29, 1999: Proceedings, volume 1597 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65895-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5101.A1 I47985 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1597.htm;
http://www.springerlink.
com/content/978-3-540-
65895-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1597.

Zinterhof:1999:PCI

[ZVU99] Peter Zinterhof, Marian Vajtersic, and Andreas Uhl, editors. Parallel computation: 4th International ACPC Conference including special tracks on parallel numerics (ParNum’99) and parallel computing in image processing, video processing, and multimedia, Salzburg, Austria, February 16–18, 1999: Proceedings, volume 1557 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CODEN LNCSD9. ISBN 3-540-65641-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5 1999. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t1557.htm;
 http://www.springerlink.
com/content/978-3-540-
65641-8;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1557.

Zhou:2009:AML

[ZW09] Zhi-Hua Zhou and Takashi Washio, editors. Advances in Machine Learning: First Asian Conference on Machine Learning, ACML 2009, Nanjing, China, November 2–4, 2009. Proceedings, volume 5828 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-05223-1 (print), 3-642-05224-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
05224-8.

Zwiers:1989:CCP

[Zwi89] J. (Job) Zwiers. Compositionality, concurrency, and partial correctness: proof theories for networks of processes and their relationship, volume 321 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1989. CODEN LNCSD9. ISBN 0-387-50845-7 (USA) (FRG). ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 272 pp. LCCN QA76.5 .Z871 1989. DM42.00. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0321.htm;
http://www.springerlink.
com/content/978-0-387-
50845-0;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=321.

Zlatev:1981:YSL

[ZWS81] Zahari Zlatev, Jerzy Waśniewski, and Kjeld Schaumburg. Y12M solution of large and sparse systems of linear algebraic equations: documentation of subroutines, volume 121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1981. CO-DEN LNCSD9. ISBN 0-387-10874-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 128 pp. LCCN QA188 .Z5 1981. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t0121.htm;
 http://www.springerlink.
com/content/978-0-387-
10874-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=121.

Ziarko:2001:RSC

[ZY01] Wojciech P. Ziarko and Yiyu Yao, editors. Rough sets and current trends in computing: Second International Conference, RSCTC 2000, Banff, Canada, October 16– 19, 2000: Revised Papers, volume 2005 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CODEN LNCSD9. ISBN 3-540-43074-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .R49 2000. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2005.htm;
http://www.springerlink.
com/content/978-3-540-
43074-2;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2005.

Zhou:2010:ACN

[ZY10] Jianying Zhou and Moti Yung, editors. Applied Cryptography and Network Security: 8th International Conference, ACNS 2010, Beijing, China, June 22– 25, 2010. Proceedings, volume 6123 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13707-5 (print), 3-642-13708-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13708-2.

Zhou:2006:ACN

[ZYB06] Jianying Zhou, Moti Yung, and Feng Bao, editors. Applied Cryptography and Network Security: 4th International Conference, ACNS 2006, Singapore, June 6–9, 2006. Proceedings, volume 3989 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34703-8 (print), 3-540-34704-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
34704-0.

Zhang:2008:PWR

[ZYBX08] Yanchun Zhang, Ge Yu, Elisa Bertino, and Guandong Xu, editors. Progress in WWW Research and Development: 10th Asia-Pacific Web Conference, APWeb 2008, Shenyang, China, April 26–28, 2008. Proceedings, volume 4976 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-78848-4 (print), 3-540-78849-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
78849-2.

Zhou:2009:DSA

[ZYDL09] Xiaofang Zhou, Haruo Yokota, Ke Deng, and Qing Liu, editors. Database Systems for Advanced Applications: 14th International Conference, DASFAA 2009, Brisbane, Australia, April 21–23, 2009. Proceedings, volume 5463 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CODEN LNCSD9. ISBN 3-642-00886-0 (print), 3-642-00887-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
00887-0.

Zhou:2003:ACN

[ZYH03] Jianying Zhou, Moti Yung, and Yongfei Han, editors. Applied Cryptography and Network Security: First Title: International Conference, ACNS 2003, Kunming, China, October 16–19, 2003: Proceedings, volume 2846 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-20208-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5102.94 .A28 2003.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t2846.

                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        20208-0;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2846;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b13996.

Zhuang:2006:AMI

[ZYRH06] Yueting Zhuang, Shi-Qiang Yang, Yong Rui, and Qinming He, editors. Advances in Multimedia Information Processing

— PCM 2006: 7th Pacific Rim Conference on Multimedia, Hangzhou, China, November 2– 4, 2006. Proceedings, volume 4261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-48766-2 (print), 3-540-48769-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        48769-2.

Zaki:2010:AKDa

[ZYRP10a] Mohammed J. Zaki, Jeffrey Xu Yu, B. Ravindran, and Vikram Pudi, editors. Advances in Knowledge Discovery and Data Mining: 14th Pacific-Asia Conference, PAKDD 2010, Hyderabad, India, June 21– 24, 2010. Proceedings. Part I, volume 6118 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13656-7 (print), 3-642-13657-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13657-3.

Zaki:2010:AKDb

[ZYRP10b] Mohammed J. Zaki, Jeffrey Xu Yu, B. Ravindran, and Vikram Pudi, editors. Advances in Knowledge Discovery and Data Mining: 14th Pacific-Asia Conference, PAKDD 2010, Hyderabad, India, June 21– 24, 2010. Proceedings. Part II, volume 6119 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-13671-0 (print), 3-642-13672-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
13672-6.

Zurada:2008:CIR

[ZYW08] Jacek M. Zurada, Gary G. Yen, and Jun Wang, editors. Computational Intelligence: Research Frontiers: IEEE World Congress on Computational Intelligence, WCCI 2008, Hong Kong, China, June 1–6, 2008, Plenary/Invited Lectures, volume 5050 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-68858-7 (print), 3-540-68860-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-540-
                                        68860-0.

Zhong:1999:MKD

[ZZ99] Ning Zhong and Lizhu Zhou, editors. Methodologies for knowledge discovery and data mining: Third Pacific-Asia Conference, PAKDD-99, Beijing, China, April 26–28, 1999: Proceedings, volume 1574 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-65866-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 P17 1999.

URL http://link.springer-
ny.com/link/service/
series/0558/tocs/t1574.
htm;   http://www.springerlink.
com/content/978-3-540-
65866-5;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=1574.

Zhang:2002:ARM

[ZZ02] Chengqi Zhang and Shichao Zhang. Association Rule Mining: Models and Algorithms, volume 2307 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2002. CO-DEN LNCSD9. ISBN 3-540-43533-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 238 pp. LCCN QA76.9.D343 Z43 2002. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2307.htm;
 http://www.springerlink.
com/content/978-3-540-
43533-4;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2307.

Zhang:2004:ISI

[ZZ04a] Kan Zhang and Yuliang Zheng, editors. Information Security: 7th International Conference, ISC 2004, Palo Alto, CA, USA, September 27–29, 2004: Proceedings, volume 3225 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-23208-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 I85 2004; QA76 A1 L43 3225. URL

                                        http://link.springer-
                                        ny.com/link/service/
                                        series/0558/tocs/t3225.
                                        htm;   http://www.springerlink.
                                        com/content/978-3-540-
                                        23208-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=3225;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b100936.

Zhang:2004:ABH

[ZZ04b] Zili Zhang and Chengqi Zhang. Agent-Based Hybrid Intelligent Systems: An Agent-Based Framework for Complex Problem Solving, volume 2938 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CO-DEN LNCSD9. ISBN 3-540-20908-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 196 pp. LCCN QA76.76.I58 Z43 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2938.htm;
 http://www.springerlink.
com/content/978-3-540-
20908-9;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2938;
 http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b95170.

Zhu:2010:ICA

[ZZLL10] Rongbo Zhu, Yanchun Zhang, Baoxiang Liu, and Chunfeng Liu, editors. Information Computing and Applications: First International Conference, ICICA 2010, Tangshan, China, October 15–18, 2010. Proceedings, volume 6377 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-16166-9 (print), 3-642-16167-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        16167-4.

Zhou:2003:WTA

[ZZO03] Xiaofang Zhou, Yanchun Zhang, and M. E. Orlowska, editors. Web technologies and applications: 5th Asia-Pacific Web Conference, AP-Web 2003, Xian, China, April 23–25, 2003: Proceedings, volume 2642 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CO-DEN LNCSD9. ISBN 3-540-02354-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 A785 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2642.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        02354-8;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2642.

Zhang:2010:EED

[ZZP^{+}10] Xiaopeng Zhang, Shaochun Zhong, Zhigeng Pan, Kevin Wong, and Ruwei Yun, editors. Entertainment for Education. Digital Techniques and Systems: 5th International Conference on E-learning and Games, Edutainment 2010, Changchun, China, August 16–18, 2010. Proceedings, volume 6249 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. CODEN LNCSD9. ISBN 3-642-14532-9 (print), 3-642-14533-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-642-
14533-9.

Zhou:2007:AMF

[ZZTG07] S. Kevin Zhou, Wenyi Zhao, Xiaoou Tang, and Shaogang Gong, editors. Analysis and Modeling of Faces and Gestures: Third International Workshop, AMFG 2007 Rio de Janeiro, Brazil, October 20, 2007 Proceedings, volume 4778 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-75689-2 (print), 3-540-75690-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.
com/content/978-3-540-
75690-3.

Zhang:2012:ISI

[ZZZL12] Yanning Zhang, Zhi-Hua Zhou, Changshui Zhang, and Ying Li, editors. Intelligent Science and Intelligent Data Engineering: Second Sino-foreign-interchange Workshop, IScIDE 2011, Xi’an, China, October 23–25, 2011, Revised Selected Papers, volume 7202 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31918-1 (print), 3-642-31919-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        31919-8.